A Complete Bibliography of the Bulletin of the Atomic Scientists of Chicago and the Bulletin

of the Atomic Scientists

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

15 April 2025 Version 2.53

Title word cross-reference

#59 [Mem80].

+ [Lin84b]. $1 [Rot96b, Dow81g, Dow82h]. $17 [Rea68]. $19.95 [Ack97]. $22.95 [Tuc97]. $27.50 [Rot97c]. $3 [Bro80b, Isa89h, Pra95]. 3 [HV18]. $3.7 [Pai04f]. $34.95 [Nor06]. $340 [Spa95a]. $4 [Sch96]. $4.1 [Rot93s]. $49.95 [Ack97]. $72 [Spa95e]. $800 [Ano96o]. $86.00 [Ano10a]. $9 [Sha00]. = [Ano92x, Lin84b]. = mc^{2} [For80]. _{2} [Bar81d, Ids81]. b^{2} [Coa88]. L + RV [Dre82a]. N [Ano60f, Gol66, Ikl60, Lak64, PK63].

-D [HV18].

.mil [Mat97b].

0 [Ano10a]. 000-pound [Ano95n]. 007’ [Dre84b]. 008 [Hib94a].

1 [Ano94u, Bat46a, Bez84, CAH84, Day76e, Hog86, LC78, Mor83c, Oom46a, Sch76a, Wal85b]. 1.5 [NA90q]. 10-megaton [NA89a]. 100th [Har79c]. 10450 [Ano55w]. 11 [Alb98c, Ano83h, Smi72a, vMC98]. 11-step [TM94b]. 115-ton [Han90d]. 12 [Ano88e]. 13 [Alb98c, vMC98]. 16 [Gui65, Har90c, Ing65, Lal65a]. 1630 [Nor64]. 16th [Ano72a]. 17 [Spa95a]. 18th [Man72]. 19 [FWCL20, Mec20d]. 1914 [Sto68a]. 1930s [Der88, Uhl95b]. 1932 [For80]. 1933 [For80]. 1942 [Str62, Wat82]. 1945 [Ano58i, Ano66s, CD48, Dow80h, Dow82h, Gro70, Ram82, Fat46, Fel81a, Sca91, Van88]. 1946 [Ano46l, Lie46, Mil48b, Rab52a, vN57, Alv87, Ham83, Nie63]. 1946/1984 [Ham83]. 1947 [Ano47v, Gre71]. 1947/1952 [Gre71]. 1948 [Eme48, Ein48a, Tel49]. 1949 [FR76, And71]. 1950 [Ano50-27, Fer62]. 1950cg [Wea79]. 1951 [Ano52h, Che52, Tur52]. 1952 [Ano52e, Ano52s, Cor53, Gre71, You76a]. 1953 [Ano53j, Ano53q, Ano55p, Bor53, Chu54b, Chu54c, Eis54c, Bou02, Win94, Zel92]. 1954 [Ano54f, Cam54, GM54, Haf54, MT55, Rab54g, SZC54, Tuf54]. 1955 [Ano55d, Ano55l, Ano55q, Ano55y, Bev56, Top56, Gre67b]. 1956 [Rab56h]. 1957 [Ano57e, BBF^{+}57, Cou57, Tel57a]. 1958 [Ano58b, Ano58d, Ano58g, Ano58l, Ano58k, Rab58b, SW59, Wei59a, Wol58a]. 1959 [Ano59b, Ano59d]. 1960 [Ano60g, Ano61i, Rab67f, Ing79b, Olu71]. 1960s [SM82, Sei60, DeB81b]. 1961 [Ano61e]. 1962 [Ano62m, NK12a, Rab63a, Rei63]. 1963 [Ano63o, Ano63q, Khr63, Pol68, Sim87a]. 1964 [Ano64h, Ano64p, Ano64q, Rab64b, Rab64c, You66c]. 1965 [Ano65d, Rab65b, Wei66a]. 1966 [Ano66a, Rab67h]. 1967 [Ano67c, Bet67, Int67, McN67a, RB67]. 1968 [Ano68b, Ano68f, Ano68g, Ano68l, Joh68, WG69, Han98a]. 1969 [Ano69b, Rab70c]. 1970 [Bar73a, Win98]. 1970s [Lea69]. 1971 [Ano71a, Jac71b, Ano78k]. 1972 [Ano72d]. 1973 [ADH^{+}73, Ano73a, Ano74i]. 1974 [Ano74a, CF74]. 1975 [Ano75b]. 1976 [Ano76a, Roc81]. 1977 [Ano77k, Gro78b, Woh78, dF79h]. 1978 [Ano78h, Ano78q, Dow81f, Mos79, Wei78e]. 1979 [Ano79b, Fel79o, MS77c, Pol79a]. 1980 [Ano80e, Ano80g, Ano80l, Ber81, Cho81, Fel93, MDB81, Shu81, Ste81a, Abb85, Mer90]. 1980s [Con90a, Dre82b, Fra83, Isa90q, Nye82, Pai80b, Sno80, vW80, Ano82i, Ber84a, Mar85a, Win90]. 1981 [Ano81d, Coh81, Fel81c, Fel81m, Isa80, Pug81, Toy82, Her83a]. 1982 [Ano82b, Ano83d, Kal82a, Pai82j, Rat83, Lyt79]. 1983 [Ano83e, Ano84f, Mor85a]. 1984 [Ano84b, Ano84c, Ano84d, Ano84k, Fel84f, Hol84a, Ham83, Hos87a, Roc85, G.84b, H.84b]. 1984/Fictions [Hos87b]. 1985 [Ano85e, Ano85v, GH85, Goo86, Roc85, Goo86, Ler80a]. 1986 [Ano86a, Ano86m, Ano86v, Hol87a, Ano86l, Ano86i, Ano88c]. 1987 [Ano87t, Ano87-45, Ano88x, Ano88y, NCA88a, NA88e, Har88c]. 1988 [Ano88f, Ano88j, Ano88m, NA88d, NA88t, NA89h, Her90a]. 1989 [Ano89h, Hol90, NA89e, NA89q, NA90e, Ano90p]. 1989-1990 [Ano91-66].

1990 [Ano90q, Ano91s, Leb91b, NA90b, NA90r, NA91f, NA98e, Kor92, La 91]. 1990s [Ano90-35, Kla90, NA89m, NA90k]. 1990s-style [Ano90-35]. 1991 [Ano91t, NA91a, NA91c, NA91d, NA91p, NA92g, SB91]. 1992 [Ano92k, NA92d, NA92p, NA93h, O’L77, Rot93a]. 1993 [Ano93o, NA93d, NA93q]. 1994 [Ano94i, Cle95, Nor94a, NA94c, NA94g]. 1995 [Ano95h, NA95b, NA95e]. 1996 [Ano96h, NA96e, NA96b]. 1997 [Ano97h, NA97d]. 1998 [Ano98i, Moo99g, NA98f]. 19th [Ano70a, Rab70c].

2 [Ano82o, Bat46b, Bro90, C^{+}87a, Gom75a, Isa89i, Isa90u, Isa91j, Kap78c, Lan90m, Nea78b, NA89i, Oom46b, Par47, Phi60, Tsi83i, Fis55]. 2.0 [VA11]. 20-year [Isa90a]. 2000’ [EE85, Asi62a, Bas80, Mei68]. 2001 [DRB92]. 20th [Ezr72]. 21st [Dan16, DM24c, Mar98c, Mec22c, Sim95, Tal18, Bul04a]. 21st-century [Tal18]. 22 [Ano94j, Isa99b]. 230 [Gol21]. 28 [Tel57a]. 29 [Ram82, SZC54]. 2s [Pik96].

3 [Ano94u, C^{+}87b, Gom75b, Mar83b, Ano90m]. 30-minute [Ark94d]. 300 [Ano68c]. 37B [Gho15b].

4 [Bet70, Bev56, C^{+}89, Hol83c, Lew65a, Sal69]. 40-Year [Win84a, Wei94c]. 4846 [Ano50w].

5 [Ano47f, Cob88, Fel80g, McD13, NBF94, Sar83]. 52s [NA92a]. 55 [Ano90-35, You56].

6 [Ark83e]. 60 [GR79]. 60-Year [Str05a].

7 [Car81]. 70 [Ano14a, Rab62a]. 70th [Mec15b].

8 [HB10, Rab69h]. 85 [Ano14a].

’90s [Isa92g]. 93-million-mile [Ano13f]. 95-year-old [Ano14c]. 96 [Hum59b]. 978 [Ano10a]. 978-3-8329-5398-0 [Ano10a].

= [Pai04f].

A-Bomb [Dow83b, Ano45c, LRD^{+}91, Jan50, Alv84, SK84, Gre64a, Shi54c]. A. [Ano45n, Ano48-38, Bea47a, Fry58, MB48, Rab63a]. A.E.C. [Ano47s, Ano48c, Kih47f]. A.T. [Ell77]. AAAS [Ano47a, Ano49l, Ano51a, Goo55, Sko68, Cal82]. Aaron [Hol83a]. Abandon [Orm85, Ear86, Squ21b]. abandoned [Gao17, NA91m]. abandoning [Cla77, Sco92]. ABCs [Rat71a]. Abdul [AH92g]. Abel [Cam89]. Ability [Ure48d]. Able [FE23, Moo94i, Wei94b]. Able-Baker-Where’s-Charlie [Moo94i, Wei94b]. ABM [Fre88, Ack87f, Ano67a, Bet69a, Cof70, Dol76, Dre84a, Ful69, Gom65, Gom67, Gro98, Hal71, Hib68, Ing68b, Isa98d, Kal68, Kre86a, Lal67c, Lal67d, Lev87, Lon68, LP85a, Lon87, Men96b, Mol69, Pai83a, Rod67, Sig88a, Sto66a, War91, Yat69, You67a, You67b, You68b]. ABMs [KP85, Rat71a]. Aboard [Jac71a, Sch89c]. Abolish [Fel55, vH79]. abolished [Sho80]. Abolition [Ano98h, Boy10, BEI^{+}55, Cra01, Gus09, Moo97f, RE63, Moo96d, Moo97c, PD89, Per20, Smi96]. aboriginal [Nel92c]. Aborted [Eas85]. Abortion [Dje72b, Dje86]. abortive [Joh86]. Above [Dav05, Cal17, Sco08b]. Abrecht [Tod80]. Abroad [Ano46-82, Ano50-35, Ano54t, Ano54u, Ano67o, Jen63, Rab68f, Sim54d, Sim54e, Wei46, Whe84b]. Abrüstung [Ano75g]. Absent [Llo71]. Abshire [LaF90a]. Absolute [For80, Yuk81]. Absolutes [Dud77b]. Abstractors [Ano79o]. Abstracts [Gar68]. absurdity [KA90]. Absurdum [She64b]. Abundance [CD77, Fre60, Pie63, Rie58]. Abuse [Fre68a, Gre77, Gre23, Isa85e]. Abuses [CS75, Klo85]. Academe [Nel74b]. academia [Kri14, Rot89b, War87]. Academic [Ano67m, Ash57, Boe57a, Boe57b, Che50, Con65, Dar57, Don57, Gla56a, Nat85a, Shu78, Vie62, Was66, GS81, Har89b, Sch75e, Din69]. Academician [Ros69]. Academies [TV71]. Academy [Ano45m, Ano46-47, Ano72f, Ano85a, Ano86w, Ano88e, BS57, Rab61h, Cav92, Sin79b, Osb58, Hym68]. Accelerating [Wil78c]. Accelerator [Ano68e, Ste77a, WG69]. Accelerators [All55l]. Accent [Smi59b]. accentuated [Gal78a]. Acceptability [Old62]. Acceptable [Fel67a, Mac77c, You47, Wei77a, Dow82a]. Accepted [Dre74]. Accepts [Tru47]. ACCESS [Ano89g, Cos02, Sch62b]. Accident [Ano80k, Ano89j, LD00, Mor09, Han90d, vH11, Sti83, And55, Mag81]. Accidental [Dav82, Leg58, Phe61, Rab58a, Pod16, vH^{+}90b]. accidents [Ano90-54, Bey80b, Fra90a, Fri11, GB11, Han90b, Kla19, Per11]. Accommodating [Bal79b]. Accommodation [Rau99]. Accomplices [Bar76b]. Accomplish [Err58]. Accomplishments [Tel54, WC17]. Accord [Bor86a, Kre87d, Bur90a, Lev88, Sak88, Sig85a]. According [Bos05, Ano75g, Ano87w, Gra81a, Isa88c, Men93b, Thi46]. Account [Bro47e, Den47, Leb91b, Woo01d, Isa90d, Kre90, McC88a]. accountability [Eps95]. accountable [BW18]. Accra [You62]. Accuracy [Pit91, Mac86c]. ACDA [Flo89b, Isa93b, Nor63, Wal89b]. Ach [Ano93r]. Acheson [Ach56, Ano46-47, CF46, Rab46a, Tel46c]. achieve [vD78]. Achievement [Chr68, Dje74]. Achilles [Ano85d, Asi63]. Acid [Car89, RC90]. Acids [Fox73]. Ackerman [Gin54]. ACM [Hau89]. acme [Hau89]. ACMs [Ver89]. Acquiring [Myr75]. Acquisition [Bin10, Sap75a, Ull63, Zah69, Mat97e]. Acquittal [Han00a]. Acronym [Ano91m, Ano91-33, Car13]. Across [Ano50i, PI92]. Act [Ano46l, Ano50-27, Ano78k, Atr06, Bun98, Dow80g, Dow80f, Dre74, Gre67d, MT55, Mil48b, Rab52a, Ric75b, Wei78e, Wil68c, Ark83c, Ark87d, Flo89b, Pro85, Ano48-34, Ano51n, Ano52m, Ano54u, Ano55b, Coh48, CH54a, DeC48, Dub50a, Lev46, Lev47, Lev48, Mar48b, McM48a, Mil52, Noy46a, Oom46b, Shi47c, Sim54e, Zei51, Zim81a]. Acting [Dys75]. actinides [Day77d, Eds76]. Action [Alb78, Ano54h, Bak75, Ber73c, Com63a, D’A70, Hor68, Hux49, Joh05, Osb47, Pri74, Sin63c, Tay58, Whe51, Wil53a, Ano77j, Ano91-28, Chu76, Day77a, Eav19, Ema21, Fri89, Gar88a, Gif13, Mec19d, Mec20c, Mec21g, Moo96m, NLS25, SV86, Sez20, Sok22, Sto17a, Sto18b, Sto19b, Tho81a, Tsi78a, Wel80, Wur91a, dTM93, Fri89, Fer53, Mye87a, Gra84b]. Actions [Ano12a, Kim87, Gur88a]. Activating [Wil73b]. Active [Mid73, Shu76, You67b, You68c]. Active-type [Shu76]. Activism [Bre93, Fil76, WWHV07, Boy84, Kuz88]. activist [Ano14c, Bes85, Moo98h, Sak98]. Activists [Der88, ML88]. Activities [Ano45c, Ano50-35, Ano54q, BAB^{+}46, DZA47, Fal68, Gal65, Tro58, Tur51, Tur52, vL48a, vL48c, Ano48u]. Activity [Kre77, Nov56, PPE22]. Acts [Ano70e, SO06, Gus11, Ano47e, Cun52]. actual [Mec20c]. actually [Cul14]. Ad [Lov74b, Tod79, She64b]. Ad-hocracy [Lov74b]. ADA [Ing46]. Adam [Ano87s, Dow82f, Mor85a, Phe56, Sco71, GC63, Ive16, JvL63, Lar63, vL63]. Adamantiades [Ano85d]. Adams [Dys72, Fel84g, Har92a, Kal68, Yud83]. adapt [Tre18]. Adaptability [Nor64]. Adaptation [Hoa69, Huq11, Fie18, MRH16]. Adapting [DC14, Mou17]. adaptive [Hil12]. Addenda [Ano83a, Ano83b, HC71]. Addendum [Ano56a, Ano68d, Ano82a, Ano83c, Ano84a, Ano17a, Duf70, Jac74a, Yut69a]. Addiction [Uhl97a, Wol91a]. Adding [Ben83]. Address [Ano46-80, Ano57h, CCD^{+}52, Eis54c, Tru53, Top56]. Addresses [Lea47, vH87]. addressing [Bru15]. Adelman [Kre90]. Adenauer [RK89a]. adequately [Van90a]. Adjusting [Nel64]. Adjustments [Mum54]. Adlai [Rab65b]. Adler [Ekl80, Sch69a]. Adler-Karlson [Sch69a]. Administered [Lev70]. Administration [Bro62a, Duf86, Lev87, Lil47c, Mar48b, Ack87g, BL79, Fie21a, Ger82, Gle21, Kin78a, Kir17, Kre86e, Mec17d, Mec21e, Mou17, Ant74, HC87, Lin84c]. Administrator [Ano49-37, Col59, Gre62, Pol66]. Administrators [Shi54d, STS^{+}49]. Admiral [Ano57c, Lap54c, Lan90k, Ano91-33, Lan90a, Lan90b, Pai82a]. ado [Ack87e]. Adrift [Ano86l]. Ads [Rot92m]. adults [Mec21a]. advance [Vad21]. advanced [CJLT24, Eav17a]. Advancement [Ano57a, Bla57, Whe84b]. Advances [Haf85, KK98a, Nix10, Eav17b]. Advancing [Was64]. Advantage [Kra73, Pik07]. Advent [Pie63, dF79f]. Adventure [DuB57b, Lun65, Dow80f, Eml66]. Adventures [Wil76b]. Adversary [Gof71]. advertised [Ano97l]. Advice [Com75a, Her88b, Low65, Mec21e, Med81, WH89, Wil78a, Ano94o, Day75c, Key81]. Adviser [Her83a]. Advising [Woo92]. Advisor [Jac74c]. Advisors [Ahe91, Eid66, Lil49a, Wei52a, Zac76, Yor76b, Yor89, Ler80a]. Advisors/Oppenheimer [Zac76]. Advisory [Ano50-36, Ano50-48, Ano52a, Ano85t, Ano85w, Ano95s, OBC^{+}49, Sap68, Ano51l, Ano52r]. advocacy [SD17]. advocate [Coh86, Ols16, Sch15b, Wei86a]. Advocates [Smi75b, Ano82c]. AEC [McE52, Ato54a, All58a, Ano48d, Ano48n, Ano48a, Ano48b, Ano48q, Ano49a, Ano49b, Ano49d, Ano49c, Ano49f, Ano49e, Ano49h, Ano49t, Ano49z, Ano49-32, Ano50a, Ano50b, Ano50o, Ano50x, Ano50-40, Ano50-47, Ano51l, Ano52a, Ano52k, Ano55a, Ano58h, Ano61k, Ano64i, Ano71g, Bos51, Bur50, Cow71, Dea53b, DMD64, Foz63b, GL48, Gre67a, Gre68a, Gre81, Hew81, Hig63, Joi54, Jac72a, Jac74b, LeR49b, Lew72c, Lil49a, Mar49f, McN48, Mil48a, Moo51, OBC^{+}49, Opp54a, Pal55, Pau70, Rab49b, RHKK78, Shi48a, SZC54, Swa49, Szi49a, Szi63a, TG70, TB70, Tru49, Tur50, Tur52, Tur53, Wil79]. Aerial [Rus70b, Fis46]. Aeroplane [WW62]. Aerospace [Sch63a]. Aether [Lit54]. Affair [AR80, Gol73b, Low66, Sim63, Ano87-41, Cha89d, Har90c, GM10, Han00a, Wei13, Wil68b, Bar67a, LaF86]. affaire [Mas90]. Affairs [Ano64q, Ano70a, Bak75, BBG^{+}85, Bol63, Gom70, GR63, Hei73, Int67, Kis60, Lon67, Opp51c, Pei69, Pol66, Rab56k, Rab56l, Rab56j, Rab62b, Rab70c, Say55, SD55, Wri48, Ano64p, Mar72a]. Affect [Gle94, Sch47, Sal16]. Affecting [Kas69, Mer24]. Affects [Lis58]. Affinities [Bix49]. Affirm [ACU^{+}54]. affirmative [Tho81a]. Affluence [Dix66, Bro79d]. Afford [Deu60]. Affordable [Alb94f, Coo12]. Affording [Ure67]. Afghan [Men92b, Zag91a, Ano94w, Ken80]. Afghanistan [A N83, Ano83a, Cha06a, Cha06b, Ege83, Jac80a, Nas03a, Nas04a, Nea80, Wit01a]. afraid [Ano90-32, Bar83b, McK25b, Sch80b]. Africa [Dow80c, Ade82, AZ91, Alb93b, Alb93f, AH93b, Alb94f, Ano86q, Aye85, Bur66, Con86, Dil66, Dow67, Dow85, Eis80, Fre77b, Flo88, For70, Gor96b, Gra85, Hec12, Hec20, IS15, Kan81, Lan85, Lar67, Lop76, Mab69, Maz69, Mti96a, Mti96b, Nai81, NF83, Oba85, Pab88, San91, Scu65, Teb85, TA94, Ung85, Was66, Yud85, Zam95d, Mun54]. African [BG14, Fel86b, Jac55, Lan70, Mag85, MM86, Rab85, And68]. Afro [Sim87a, Mah59a]. After [AG63, Ahm00b, Ano50m, Bec75, BV97, Blo71, Bou62, Cas04, DS49, Dro24a, Ein79a, Fel68a, Fel82a, Fer68, FvH99, FC88, Ger67, Gri83a, Jan59, JS97, Kal88, Kap72, Kra57, Lew69a, Loc92a, Mar89d, Mel71, Mor50b, Mul69, Nat88, NA98a, NA98b, O’K73, Pan88, Pra22, Pri58b, Rab47m, Rab51b, Rab52c, Rab57b, Rab57c, Rab63b, Rab73a, Rom82, Rot88, Rot01a, Sak88, Sas17, Shi54a, Shr70, Tho88, Tsi89c, War59, War88, Wea82, Wie48, Woo97c, Zub22, Abr84, Ano69c, Ano69d, Ano90-54, Ano94j, Ano98m, Cla77, Cra87, Dea83, E^{+}84, Fis90, Fra90a, Gal81a, HBD17, Ing63a, Isa90c, Kis94, Loc14, Mar98a, NF13, NA90k, Nor94b, Pod24, Rab20, RKK22, Sch84a, Sch16, Sea85a, Set16, Shi84, Sim16, Sim88, Sim91, Smi83, Smi85d, Sok22, Som91, Tsi83b]. after [Woo97e, vH80, Shi84, Mul65, Wag61]. after-action [Sok22]. Aftereffects [McC96]. afterglow [Isa92i]. Aftermath [Mor09, Rab75, Ito14, Mec17a, Sto16]. Afterthoughts [Rab51a]. Afterwar [Bri86]. Again [Alb64, Ama03, Ano67e, AK01, Mac57, Mar64i, Rot02f, Sig01, Wei04, Ano90-48, Ano93y, Ano94k, Cha78b, Cla91, Dro24f, Els90, Han79, Isa87c, Isa96f, Knu89, NA94e, Sha97, WG82, Woo98c]. Against [Ano54a, Ano62p, Ano69a, Bar69, Bon51, Cap98, Cha70b, Dys64, Gil83, Hal71, Keu83, KASN52, Klo85, Lac69, Law69, Lov73, Lyn69, Mar58b, McG67, Mor85a, Mul53, Nel74b, Olu71, Pet81, Sil64b, Ste76a, AP88, Ano77j, Ano86i, BG14, Bud22, Esp14, Fra75, Hay55, Hun14, KW86, KSG15, LB78a, Mia15a, Nic54b, Osb82, RS88a, RK88, Sim60c, Tan83, Tom14, Wat19, Mac84b]. Aganbegyan [Cam89]. Age [Ano54-30, Ano58l, Ano70d, Ano10a, Bel93, Cus48, Ded70, Dje74, Don89, Dow82a, Gla58, Gre67d, GR63, Han98a, Hei70, Hei64, Hel07, Her90b, Ima70, Les63, MF10, Mil88, O’C87, Rab69i, Sac84, Saw81a, Sch59b, Smi70a, Ste77b, Thi59, dF79d, AM86a, All15b, All17, Ano81k, Ano11c, Ben76a, Ben76b, But97a, Doy17, Dum76, Gar93, GN10, Jon84, Lif85, Man84a, Mec15a, Mec16d, Mec23c, NB77, Rin84d, Ros89a, Wil77b, Air55, All53, Ano50c, Ano73d, BW59, Bet62a, Bri86, Cal84a, De 70b, Dew48, Dow83a, Fin90, Hol66, Hus53, Kla99, McL53, Mic69, Mor62c, Pei70, Pri68, Rab59j, Rab60a, Rab72f, Rab72g, Rab73f, Rih65, Roc81, Sch63a, Sem59, Ste76a, Sto59b, Tel49, Wea85, dJ64, Bou61, Bro60a, Cow65, Cro65b, Dow81b, For86, Gen86, Gol95a]. Age [Gol96a, Got87, Gus83, Gus85, Hol86b, Lan56a, She60, Sto58, Gue80]. Age-Old [Ded70]. Aged [Ano94m]. Agencies [Ano46-66, Pow47, Ric75a, Sta96]. Agency [Ano78g, Cla74, Dem60, DZA47, Hoe88, Lal86, MEA^{+}47, Pri69, dRVT^{+}47, Sub70, Smi75a, dRMT^{+}47, CD80, Eps77c, Kre88b, Bec58, Col59, Jos61, Lil47c, Loe84, Opp47b, Ran59, Smi84b, Voh71a, Zlo47, Ano78n, Kin80b]. Agenda [Bar84, Bin84, Car84a, Coa84, Cra84a, Day75d, Fel84a, Isa00c, Isa01e, Kin84a, Kor92, Man84c, Tru84, Wal85a, Ano86x, Ano91n, GK21, Hib94e, Isa91n, LL16, Per91b, Sim90a, Sim90b, DRB92]. agendas [Fer88]. agent [Ano91-36, Ano95j, Hib94a]. Agents [Ang60, Lwo60, Ste07a, Mei77b]. Ages [Rot05a, Mon84, Fel76i]. Aggression [Sei73, Har88d, Tho80g, She89]. Aggressive [Moo01c]. Aggressor [Thi53]. aggressors [CMM95]. Aging [Str07a, RE19]. Agnes [Ada73]. Ago [Ste65a, Bul02a, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano78a, Ano78b, Ano89q, Ano89z, Ano89y, Ano90-56, Ano90-57, Ano90-62, Ano90e, Ano90f, Ano90h, Ano90c, Ano90g, Ano90b, Ano90d, Ano91e, Ano91f, Ano91d, Ano91b, Ano91g, Ano91h, Ano91c, Ano91a, Ano91i, Ano91l, Ano92i, Ano92g, Ano92j, Ano92c, Ano92d, Ano92a, Ano92b, Ano92e, Ano92f, Ano92h, Ano93a, Ano93e, Ano93f, Ano93d, Ano93h, Ano93b, Ano93i, Ano93g, Ano93c, Ano93j, Ano94g, Ano94c, Ano94d, Ano94a, Ano94e, Ano94b, Ano94f, Ano95e, Ano95f, Ano95b, Ano95g, Ano95d, Ano95c, Ano96-28, Ano96c, Ano96a, Ano96e, Ano96d, Ano96b, Ano97e, Ano97b, Ano97f, Ano97a, Ano97c, Ano97d, Ano98b, Ano98e, Ano98c, Ano98a]. ago [Ano98f, Ano98d, Gom82b]. Agony [Ano99a, aI98, Isa90c]. Agrarian [Gra66]. agree [Ano91-32, Cra96, Fel80e]. Agreeing [GA08]. Agreement [Fed59a, Fel75b, Lon75, New54, Sha60, Ano86w, Ano94m, Arn47, Bon83, Kap90b, Kre87c, Nor11, Par17, Rot81a]. Agreements [Ano63-28, Rat72, Ano89c, Bil85, Dom82, HA85, Win84b]. Agricultural [Ano64a, CPV01, Mab69, Olu66, Yud85]. Agriculture [Ano66d, Joh64a, Joh64b, Joh83b, Rob66, Sax49, Sch68h, Wal50, Mat18, Teb85, Wun82]. Agro [Vof72]. Agro-Industrial [Vof72]. Agroforestry [Spu80]. Ah [Rot03a].

Ahead [Cra70b, Hol05a, Isa95f, Jac74b, Pri77, Sea67, Sei49, Ack90g, Ano13b, Dro23a, HFMH89, WL84, Ano84a, Dow84a]. Ahmed [Lal67b]. AI [CW23, Dem19, Dro23d, Gau24, Gei16a, HD23, HK21, Mec23c, Rof19, Sch24, Shr23, Sto23, YWS20, YSW21]. Aid [Ano55c, Ano62-28, Bar81i, Bow57, Fis46, Han02, Loc93a, Pre60, Rab67d, Rev68, Sko67, Spu72, Gra85, Isa93d, Kor92, Kri86, Lei90, Lyn86, McN77, Mos88, Sch83, Sol95, Sol90, Col84a]. aids [Joh94b]. Aikenhead [Dow80g]. Ails [Sch68h]. Aim [Cle99, Goo03, Gro05, Isa02e, Sch74c, Pit91, Rec89]. Aiming [Isa98d, RC87]. Aimless [Rab68e]. Aims [Boy00, Eme47, Lon85b, Lov63, Nit57, Rab60a]. ain’t [Dun90]. Air [Ano73h, Axe51k, Bai92, Ber70a, Bet46, CG04a, CR71, Dix65, Fer69b, Flo88, Gal56, Gol61b, GF73, Jer87, Kur79, Mor84c, Pot53a, Pot53b, She83a, Ano93u, Ano94n, Bet20, Col94, HNB81, NA88b, Ano92m, HWG95, Joh94a, Mar65d, McM54, Wal55]. Air-breathing [Mor84c]. Air-Nuclear [Gal56]. Airborne [Pri03, Flo89a]. Aircraft [Ano93k, Haf49, Ste95, NA87l, NA88p, Ano88d]. Aires [Gol90d, Lem88]. airport [Sch89a]. Akademgorodok [Lei72a]. Aken [Sin67a]. Al [AH92k, Nas03b, Dro16b]. al-Faisal [Dro16b]. Alamogordo [Rab69d]. Alamos [Ano45n, Ano22c, Bra54, Eis54a, Fel81a, Fer54a, Gus11, Man49a, Man49b, Tay98, Ano96m, Cor95, Gus99, Hun04, McD74a, Mit70, Sai09, Smi70a, SW80b, Tay98, Tel54, Wei83c, Wei85c, Wil75f]. Alan [Ano78d, Ano87n, Div70, Div79, Gou78b, Kal62b, Lep89, dF79h, Ano15a, Mec21a]. alarm [KP85, Sta93, Yor75b]. Alarming [Fol74]. alarms [NA90n]. Alas [Isa89m]. Alaska [Ano89r, Mus70, NA89f, O’N89a]. Alastair [Kri85, Sei60]. Albania [Woo02, Woo05]. Albert [For80, Hen71a, Sha54b, Wal85a, Bro79c, For80, Opp79a, RF55, RF79, Sea79, VFIS63, Woh78]. Albright [Rot93a]. alchemy [Gib91a]. ALCMs [NA92a, Ver89]. Alcock [Lyt79]. Aldermaston [MW98]. Aldous [Doh58, PMRP50]. Aleksandr [Red69, Gut93a]. alert [Oko82]. alerting [vH97a]. alerts [NA90m]. Alex [Eav17c]. Alexander [A.54, Ing56c, Nei82a, Zim81b, Ano87o, Ano96f]. Alfonso [Bar67b, Fel82g]. Alfred [A.52, Ber79, McC93, Mul72, Nor64, Sto58, Ano84b]. Alfven [Ano71d, Jac73e]. Algebra [Hib68]. Alger [Dow81i]. Algeria [AH01]. Alice [Hig66, AK93, O’T00]. Alice-in-Wonderland [AK93]. Aligned [Jac63b, Sal75a, Eps96b]. Aligning [Bul08a]. Alike [Rot02a]. Alive [Jos61, Eps91]. Alkaabi [Mec16f]. all-nuclear [Lap76]. all-volunteer [Jan81b, MJ80]. Allan [Abb93, Fri72, McG84]. Allaying [Ing59a]. Allegations [vCM99a]. Allegiance [Dug91]. allegory [Bar77d]. Allen [Ano78n, Kli90, Mor62b, Pla57b, Val73, Foe07, Gol88a, Jac73d]. allergies [Ano85n]. Allergy [Bha03, And85, AF85, Fry85, Gri85, Mor85c]. Alliance [Mot48, Mul66, Som91, Sta85, Sul58a, CAM20, Bar84, Ken66, Sha82, Swe86, You63b, Cah76, Mul65, RK93, Stu89, Why62, Wil84, You66b]. Alliances [Pow90a]. Allied [Gro86]. Allies [Sim54d, Ber93b, Jac82d, NPS90]. Allison [Wal85a, Wei66a, All48, Ano10b, Ano12g, Mec22b]. Allocating [Kay66, Kis66]. Allocation [Kon67]. Allowed [Jac62a]. allure [Hec19]. ally [Gel90]. Allyn [Leb94]. Almoners [Wie47b]. Almost [Ack99, Ano46-100, Ber76b, Cle04b, Day75b, McK25b]. alms [Rot93c]. Aloha [Ano91-33, Ano94k]. Alone [Asp80, Cha76, Bul02c, Ter04, Rat92, Han75]. Along [Mos92, Fer10, Zim55]. Alperovitz [Hig66]. Alpha [Moo95a]. ALPS [CLD88, Sch88a, Gom75c, Hus24]. already [Gei16a, Krz20, McK23b, Pri80]. Also [CMM95, Gou75b, Kat75]. Alsos [GP92a, Mor47a]. Altering [All18]. Alternative [Ano78n, BAA^{+}75, Bin10, FT76, Gir75, HKC78, Sch78a, Som84b, Ure47, AB77, Ano12e, Bro76b, Her79, Por19, Sha22, Mul65]. Alternatives [Ano12e, Bra84, Fer60, Fie78, FK71, Gre63a, Leo64, Mar78a, Mul65, Sim55, You67a, KB89, Som84c]. Altruism [Hul78]. alumni [Ano95j]. Alva [Ano86a, Fel82g]. Alvin [Moo96c, Mot68]. Always [Ano94n, Dow80e, Eps75c, Lan89e, Leo19, Mak95a]. Alzheimer [RC90]. Am [Gou74]. Amann [Bei84]. Amarillo [Gre87]. Amazon [Dro25a, Fly03]. ambassador [Mec16c, Mec16f, Sto18a]. Ambiguity [Pan71, CF87, Par17]. Ambiguous [Ano72e, Int92, Lew05]. Ambitions [Ano91s, Ano91-66, Ima93a, Per02, SS90b, Fie24]. ambitious [GK21]. Ambivalence [Mar62]. Ambivalent [Fel78c]. Amend [Ano55b, Loe87, Rab52a, Tho87]. amended [Rab52a]. Amendment [Rab47j, Tel46c, DG79, NK86b, Wee95, Kal89, Kal53, Mel53b, Mel53a, Pow86]. Amendments [Ano46-91, CH54a]. America [Abb85, Ada87, And78a, Bla85a, Boy87, Cra84c, DeB81b, Fay75, Gan84, Gom71, Ham93, Hig66, Mar85a, Moo95b, Moo95e, Mor82a, Mos92, Mul65, Mye87a, Rau91, Rod70, Ros90a, Spi75, Swe88, Uhl97a, Uhl98a, Van03, Mos84, Wio84, Alb89a, Ano60q, Ano65e, Ano99b, Ano03h, Ano12d, Bar66b, Bon72, Boy84, Bro72, Bur74, Cli77, Coa84, Dea95, Den14, Dil66, Eav17a, Ekl80, Fal70, Fis61a, Ger17, GS81, Gol82a, Gom83, Gre73c, GM14, Gus04, Har86a, Hor75, Isa99c, Joh79, Ken78, Ken83, Kla77, Kur79, La 83, Lew73b, Lop66, MK02a, Meh21, Mil69b, Myr72, Por73, RL14, S.80, Shi52, Shi87, Smi65, Smi70b, SL83, Soa96, Spi71a, Ste89b, Str81, TS22, Ure50, Wal69, Wol80, Wol17, Woo00c, Bar67b, Der88, Dow80j, Dow83d, Gol92a, LP92, Mar79b, Rab71b]. America [Sla88, Wil69c, Woo65, Gen84b, Lap69, Mor87c, Mos90, Sei81]. American [Abb93, ADH^{+}73, Ano46-27, Ano46-35, Ano46-83, Ano46-89, Ano48u, Ano51i, GS51, Ano52y, Ano69k, Ano84h, Ano98g, Bet75b, Bri89, Cro65b, Dow80a, Fla80b, For86, Fra61a, Fra91, Fre95, Gal86, Gol69b, Gro50, Gru70b, Gus83, Har85b, Har86e, Her83a, Her85, Hig48b, How65b, Jer87, Kel79, Lyn69, Mar89e, Mar98c, Mil69a, Moo97b, Mor85a, Mor62d, Mos81b, NA92b, Olu71, Ost72, Pal64b, PMRP50, Rab69b, Rab73d, Ran85, Roc83a, Sco71, Sei60, She83b, Sim87a, Sim54a, Sin67a, Smi54a, Tod78b, Uhl95a, Val86, Win90, dF79h, AC02, Alb54, Ano45m, Ano46b, Ano46o, Ano46-88, Ano47e, Ano47w, Ano48e, Ano50-36, Ano53n, Ano54b, Ano54j, Ano54t, Ano57a, Ano64j, Ano65g, Ano67m, Ano72f, Ano85a, Ano91-47, Ant78, Aro51, Aro52, Ast67, Aue03a, Bar46a, BAB^{+}46, Bet63, BS05]. American [Ble54, Blu23, Bra49, Bro82, Bro72, Com81, CYP14, Chu76, Com74, Cor58, Doy53, Dun49, Ebe47, Ebe20, Edg52, EBU^{+}52, Fed55, Fei60, Fei61, Fel68d, Fel77o, Gil18, Gla56b, Gom75a, Hai54, HKL^{+}49, Hek60, Het55, Jen63, Joh04, JC54, KW15, Kal70, Kat54b, Kis72, Lal67f, Lap48, Leb79, Lil47a, Mas48a, Mer54a, Mit62, Nat55, Opp51b, Opp53, Opp63a, Osn77, Pre58, Pal77a, Pau57, Pip55, Pon17, Rab46f, Rab54c, Rad53, Rid47, Rum54, S.46, Sag93, San66a, Sax49, She69, Shi47a, Shu78, Sim54d, Smi72a, Smi87c, Szi61a, Tod55, TM97, Tsi83f, Tuc66, Wal82, Was64, Wei55a, WW80, Wri49, Hyd75, Lan56b, Dro23k, Pol93, Row88, Sco74, Wig62b]. Americana [Szi60a]. Americanization [Ano94k, Hor68]. Americans [Mes84a, Ano90-52, Cra84g, HSV19, Pol93, Zwe84]. Americas [Lin73, Sea73]. Amerika [Ano65e, Ano90-30, Rot93b]. Amid [MFN17, Rot01c, Jor18]. Amino [Fox73]. Amitai [Cha62, Cow65, O’B75]. amity [Sha81c]. ammonia [Mac22]. Amnesty [Ano83d]. AMOC [Dro25b]. Among [LT70, Men69, Mil70, Swa66, Aug56b, Con81b, Dje76]. Amory [Nef78, Nel84]. amplification [Kas12]. Amrine [Mar59a]. Amundson [Boy05]. anachronisms [Bar92a]. Analogy [Fas66, How63, Lev76, Smi69b]. Analyses [Ano85c]. Analysis [Bat46b, Ber77b, Cra83b, Dav76a, Ler46, Lev46, Mor49, Sco70, Shi51, Tur73, Wea75, Woo82, Fri11, Ise92, Osn77, PPE22, Ric76, Wil68c, Hol83a, Kat53a, Mar89e]. Analyze [Fei60, Fei61]. Analyzed [Gou47a]. Analyzing [Pol67]. Anatol [Bre65]. Anatoly [Cra88]. Anatomy [Ano78k]. ANC [AH93b, Bat98]. ancestor [Met24]. Ancient [Nel93a, Fre77b]. Anders [And72]. Anderson [Dow83f, Nie63, Kli81]. André [How65a, Mas66]. Andrei [Bou87, Hol91, Bro79c, Hol90]. Andreopoulos [Lyn78]. Andrew [Bam85, Bar73a, Dow83a, Gan82, Han94, Lan98, Lei92, Nor06, Sil71]. Andrews [dF79e]. Andrey [McC88a]. Andrzej [Sar78]. Andy [Rot96b, Lan90d, Smi90c]. aneous [Loc91b]. Angel [Hul76]. Angela [Tic94]. Angeles [Coo74, Ano73h, Mar73]. Angelopoulos [Sto58]. Angelos [Sto58]. angle [Arb19]. Angles [Ano96l, Moo97p]. Anglo [Tho53]. Angola [Ven99a]. Angry [Dun74, Gou63]. Angst [Pea99]. Anguish [Rap65b]. Animal [Fra72, Jac74b, Kap66]. Animals [Lan56b, Tho80g]. animosity [Bal84b, Mac84a]. Anita [Lyt79]. Ann [Ano81l, Gol92a, Gol93a, Kre90, Sei71b, Sei73, Tay98]. Anne [Eck78, Har81, Sch54a, She79, Dro21b]. Annex [Ano60a, Ano60c, Ano60d, Ano60b]. annihilation [Aus81, Gus83, Kal63a]. anniversaries [Kis82]. Anniversary [Ano48-27, OJCP^{+}47, Rab56k, Rab56l, Rab56j, Ure45, Wei83c]. Announced [Ano46n, Ano46-63, Ano47-29]. Announcement [Ano49g, Ano49-34, Ano49-38, Ano52b, Ano90j, Ano92n, Ano93l, Ano93m, Ano53o, Ano84k]. Announcements [Ano48n, Ano86b, Ano86c, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88a, Ano88b, Ano89a, Ano89b, Ano90k, Ano90l, Ano91o]. Announces [Ano46k, Ano49a, Ano49p, Kal83c]. Announcing [Bul04b]. Annual [Ano48-28, Ano48-29, Ano49k, Ano51y, Ano90m, Bar79a, Bar80b, Bar81a, Gar68, Tod79, Cra83a]. Anoushiravan [Jon90b]. Anshen [Fer53]. Answer [Adl57, Bir76, Dav57a, Koc57, KFT61, McC56, Nie55, Rei67, Rus58a, Wyl57, HSV19, Pau77, ZRG83, Ing56c]. Answering [DMJ98]. Answers [Ano47w, Chi59a, Gal90a, Gal90b, Lal64j, Tru48a, vH58, Jon13, Lee91a]. Antarctic [Dro20a, Eml66, Fuc70, Cra70a, Dan70, Fuc70, Jon70, Lew63b, Lew70a, Lla70, McE70, Pot70, Rub70, Smi69b]. Antarctica [Cra64, Ano70c, Cra60, Dou68b, Fed70, Gou70, Hel70, Mer24, dQR70, Sha84c, Sie24, Smi70c, Sol70, WM78, Woo99a, ZSA73d]. antecedent [Gol88c]. Anthology [BB63]. Anthony [Kal78, Mei68, Pon67a, Ray81, Wyl56, Dro21c]. Anthrax [BST02, Gre83b, Zil83a, Zil83b, Ber87b, JL87, Ano83a]. Anthropocene [WSG^{+}15]. Anthropological [All68, Bat46a, Bat46b]. Anthropologist [Ber64]. Anthropology [Man84a]. Anti [Ano70g, Ano74e, Ano98n, Bal86a, Bet69b, Cha69c, Dys69d, Ful69, Jac74d, Kam71, Kri95, Lov73, Nel81, Por19, Rat69, Ros75, Ste69f, Tur76, Wit01a, Yor69, Buc81, Chi82, Dav79a, Fie24, Gor76, Mag16, Pom77, Wit01c, Ano86d, BCB81, MD81]. Anti-Famine [Cha69c]. anti-hero [Gor76]. Anti-Israel [Tur76]. anti-MAD [Buc81, Chi82, BCB81]. Anti-Matter [Ano74e]. Anti-Missiles [Bet69b, Dys69d, Ful69, Rat69, Ste69f, Yor69]. Anti-Nuclear [Lov73, Nel81, Dav79a, Mag16, MD81]. Anti-nuke [Ano98n]. Anti-Pollution [Ano70g, Jac74d]. Anti-Poverty [Kam71]. Anti-satellite [Por19]. Anti-SDI [Ano86d]. Anti-Serb [Kri95]. Anti-Star [Bal86a]. Anti-Submarine [Ros75]. Anti-Taliban [Wit01a, Wit01c]. anti-test-ban [Pom77]. anti-vax [Fie24]. Antiballistic [Kal68, RA67]. Anticipation [Smi59b]. Anticipations [Mum54]. antidrug [Wal90e]. Antimatter [Ste74]. antimicrobial [Pop19]. Antimissile [Tal65, LP10]. Antinuclear [Swe86, Wal83b, Ano14c, Cle87a]. Antipollution [Net69]. Antisatellite [Sig85a]. Antisubmarine [Ano87q]. antitank [Ano90-37]. Antonio [Dow83f]. Anxiety [Abb93, Ano93-30, BF05b, Kia93, Mul69, Wor06, BDF^{+}92, vV23]. Any [Bro81, Gil77, Pot70, Ano90-30, Ano92t, Hor88c, Nor82, Pot97, AK98a]. Anyone [Kob70, Lor92, Oze92, Isa89m, Lei89a, Wal89c]. Anything [Bou74, Sch59b]. Anyway [Dow83e, AH91n, Ano91-33, Ano92q, Tho88]. Anywhere [Sør80]. apart [NA93s, vH89]. apartheid [Ack90b, Dav93]. Apathy [Lev53, Boy84]. Ape [PMRP50]. Apes [Kil78]. Apocalypse [Kee07, Moo93c, SS24, Ack91c, Cas19, Sto18c]. Apocalyptic [Tod64a]. Apollo [Fol75, Ano72e, Bro67a, Fox73, Kui73, Lea69, Lew68a, Lew68d, Lew71c, Lon69a, O’K73, Rab69d, Rab69h, Spa69, Smi72a, SS73, Ste69b, Stu69a, vB69]. Apologia [Shi49, Pat78c]. Apologies [Ano53-36, Ano78c]. Apology [Min94]. app [Ban22, FJ12]. Appeal [Ano58a, Ano67f, Ano69a, Ano73b, BEI^{+}55, Con46a, Day74b, Gab86, Hay55, Hux49, Int50, Pau57, RE63, Sch57b, Ano79h, Ano80o, Ano82t, Cas84, Kas75, Lev78, Uni78, You81, Wri47a]. appeals [Fel82l]. Appendix [Ano60e, Mar78a, Ano86e, Mah59a, NAB99a]. Application [Aeb48, Hei47, Joh60b, Kam46, Moo00c, Rau69, Sab72, Spo50, Wei70b, Wri48, Tel47b]. Applications [Ano71h, Bla69a, DuB69a, Fre72a, Kis66, Sag86, Wal69, Jas81]. Applied [Ano46c, Ber72a, Dew55, Tel66a, SW59]. Apply [Mas57, Ano79k, Ano96n]. Applying [Gaf65]. Appointed [Ano46p, Ano46-43, Ano46-66, Ano52r, Boa91]. Appointment [Ing55b, Lam69]. Appointments [Cou54]. Appoints [Ano46-99, Ano56n]. Appraisal [Gro53c, Lal67h, Lon57, Mul65, New64, Ore60a, Rat72, She83b, Wel60, dR49, Kri85, Don86b]. Appraise [Ano62-27]. Appraises [Ano49-40, Bet60]. Appreciation [Mus78, Rot95b, Tur54]. Approach [Bat46a, Bat46b, Bow57, Cus48, Gus83, HFO63, KK72, Kno66, Leg57, Myr75, Ols63, Ore61a, Pla53, Sac72a, Sel69a, Spu72, Ura01, Zar64, dN71, Ban11, Cas77e, Dan16, Gor84, Got20, Hol86d, Kir17, Kro15, Mec17d, Mia15b, Mor77, Pai83e, RG13, Sol79, Yin15, Nor64, Haw47]. Approaches [Dos72, Lal65b, Lee73, Arn22]. Approaching [Spe18]. Appropriate [Ano81f, Dow80e, Gar88a, Row77]. April [Ano47v, Ano58g, Ano80a, Ano81a, BBF^{+}57, Eme48, McC54c]. Arab [Al 11, Ker68, Al 15, SW07, Zah69]. Arabia [Dro16b, Sez16, Sez20]. Aramco [Sez16]. Aranha [Kih47f]. Arava [Vof72]. Arbiters [Bar63]. Arbor [Ano81l]. Archaeology [Haw73]. Archaic [Fre72a]. Archer [FE23]. Archibald [Ano68g]. Archie [Uhl97b]. Architects [Aur72d]. architecture [Bru82, Cra82d, Nor11]. Archives [For80]. Arctic [Bac12, Bli70, Bor69, Dro19, Gus69, Hue12, Krz20, Mor12, RV64, Wal08]. arduous [NH08b]. Area [Ano73h, Bak75, Din78, Rot98a]. Areas [A.51, Bar60, Bar61a, Foz61i, Ran55a, Rei63, Ano93w, Mar54d, McE88]. Arena [Pei69, Day75c]. Arenas [Tho75]. aren’t [Ark02b, Car13, Sch05e]. Argentina [Alb89c, Est01, Gol90c, Gol91c, Kra92, Lev77a]. Argentine [Esp14, Tel66b]. Argonne [Ano46-73, Dan48]. Argument [Ago77, Bar69, Chi78, KD59, McN63a, Ols16, Osb82, Pan84]. Arguments [Gil83, Bre82, Bro90, HGF17, Kre91b]. Ariane [Bas80]. Arid [Vof72, WC73]. Arise [Ano56g, Wal74]. Arising [Com57a]. Arizona [Lew64a]. Arjun [Pre96]. ark [Mye82b]. Arkin [Cal84a, Lev84b]. Arm [Ada87, Ric79]. Armageddon [Bri86, Dow83d, HHH07, ML09, Rab48g, Sch05b, Don86a, Mar97a, Rod85]. Armament [Axe51k, How62, Ing56b, Leo64, Ano86o, Bar76d]. Armaments [Ano47q, Ano86l, Bat46a, Bat46b, Fel72a, G.84b, Pol61, vW80, Ste77b]. Armand [Wei86b]. Armed [Cra83a, Cra83b, Fin65, Gol91b, Har92b, MW75, Roc81, Val97a, Ano93-32, Cot19, Kla91, McD13, Tsi77a, WF16, Wie47b]. Armenia [Isa94h, Var96a, Var96b]. Armies [Axe51k, LP92, For92, For20]. Arming [Har01, NA89f, TM94c, Men95]. Armistice [Ash60, DB50]. armor [Que81]. Arms [Ali57, And56, Ano52-27, Ano54t, Ano54u, Ano54v, Ano60f, Ano61i, Ano66l, Ano67b, Ano71b, Ano82c, Ano85c, Ano86k, Ano87i, Ano87n, Ano87-39, Ano89c, Ano89g, Ano89e, Ano90r, Ano90p, Ano91s, Ano91r, Ano95l, Arb19, Ark87a, Ark87b, Bal89b, Bar81c, Bar87, Bar85c, Bat98, Bec71, Ben01a, BPG89, Ber89b, Ble89, BGR61, Bor82, Bot61, Bow89, Boy89a, Bra86a, Bra86b, Bun89, But04, Cou57, CLR74, Cah75, Cav55, Cav56, Cav61, Cir07a, Cla74, Cla81, Cle67, Cof70, Cra84d, Dan16, Dav69, Day76a, DeW62, Deu83b, Deu84a, Deu67, Div79, Don89, Dow80e, Dre89a, Ein50, Eps77a, Etz63, Evr78, Fed63, Fei92, Fel61, Fel70b, Fel70a, Fel72a, Fel75a, Fel75b, Fel78a, Fen91, Fer60, Fie87, Fra61b, Fre84a, Fry65, Gab01, Gad89, Gan82, Gar89, Gar81, Gol95d, Gom61a]. Arms [Goo86, Gra73, Gra89, Gra84b, Gre62, Gur75, Gus83, H.84b, Hal89, Hal75b, Han87a, Han96a, Han98a, Har86e, Har91b, Hel77a, Hen59, Her85, Hum72, Ing56a, Ing57a, Isa87a, Isa91a, Isa96c, Ise92, Jac74b, Joh05, Kal78, Kin80b, Kin84a, Kin86, KB89, Kre85a, Kre88d, Kre89a, Kri92a, Kru74, Kum04, Lal64d, Lal65f, Lal66a, Lal66c, Lap72, Lei72b, Lep89, Lev62, Lev89, LC09, Lin84c, LW61, Loc93c, Loe84, Lon85b, Lut89, Man89, Mcc87, Men86, Mey54, Mic61, Mor87b, Nit89, Nye89, Pai82c, Pan71, Pan86b, PMRP50, Pos87, Rab61a, Raj10, Ras63, Rat71b, Ren99, Rep85a, RK89a, RR88, Rob91b, Rod67, Rot91b, Rot93c, Rui68, Sch65a, Sch89a, Sch63d, Sch71b, SD86, Sco70, Sco72, Sco74, Sea89a, Sea69, Sha81b, She79, She89, Sie06b, Sil64a, Sin05b, Slo68]. Arms [Smi85c, Smi89, Soh61, Ste95, Sto64, Sto65a, Str72, Tee65, Tel82, Tsi72a, Uhl96, Uhl98a, Van89, Wal89a, War78, War89, Wes74a, Wie67, Wie70, Win84b, Wit01b, Yor69, Yor70, Yor73, You65a, You65b, You66a, You66c, You76b, Yut69b, Zhe90c, dF79h, vM89, vW57, Ack90i, AH20, APR17, Ano87v, Ano87-30, Ano90-52, Ano95n, Ano11f, Arb16, Ark84b, BL79, Bel20b, Bre82, Bug22, Cha87, Day77a, DeW84, Dea87, Dow78, Dow82h, Dre82a, Eis90, Eps92b, Fal87, Fer10, FKM87, For91, Fra76b, Gal81a, Gar84a, Gei16a, Ger13, GK21, Gor95a, Gor84, GK85, GJ84, Hal83, Har84a, Har87c, Har87d, Har87e, Har90d, Har93, Hes85, Hin18, Hol85b, Hus90, Ibr13, Isa81a, Isa85a, Isa85d, Isa86a, Isa86c, Isa87b, Isa87c, Isa88b, Isa88c, IG91, Isa91k, Isa92j, Isa95e, Jac83, Kin78b]. arms [Kis79a, Kis79b, Kla76, Kla77, Kla84, Kla85, Kla97a, Kor19, Kre86b, Kre87e, Kre90, Küh17, KKS22, Lam87, Lat84, Lei78, LH91, Mac91, Mac81b, Mac86c, Mat85, McM05, Mcc88c, Mec20b, Mec20g, Men96b, Mil81, Mor83c, Oel12, Pai85e, Pai85c, Pan81, Pie80, Pra89c, Rof19, RKK22, Sal80b, Sch77a, Sco77, She85c, She85d, She86b, Smi85a, SS88, Ste77b, Ste77c, Tro20, Tsi83c, Tsi84b, Tsi89c, Tuc10b, Uni78, Vad21, Veg97, You78a, Zal19, Zer13, Ahm03a, Ano86i, Ano86s, Bet69a, BGR61, Dow81d, Fei92, Kin80a, Smi84b, Ano89d, Bra87, Bul67, Kre93, Lap61a, Sch79b, Siv76, Ste86c, You68a, dF79a]. arms-control [Mac86c]. Arms-Race [Gar81]. Army [Khr96b, McE88, Pow01, Sim88, Arn98b, MJ80, Rec89, Rot96f, Str94, Ano77n, Ano80i, Ano92p, Aur46, BS47, Bro47a, Fel80k, Fre93b, Gou47a, McM54, Pai80a, Dow81h]. Arnold [Bou02, Bro57a, Lyt79, Pow91b, Sco79, dF79f, You76a]. Arnove [Whe84b]. Aron [She61, Sei62]. aroused [Bar82b]. arrangement [Ano95j]. Arrangements [Lal65c]. arrived [Krz20, Zit17]. Arrogance [Gol67a, Nac69]. Arrow [Sch75d]. arrows [Plo89]. Arsenal [ACH84, Lew05, Mie74, Tra90, vH97b, Ano87-35, PN09, Zha12, Cal84a].

Arsenals [Ark84a, Bla82b, Rob75b, Str07a, Bar78b, Bar79a, Bar80b, Bar81a, Bet89a, FUvH85, Kas77]. Art [Ben59, Bro13, Cra83b, Jou52, LS59, Mil86b, Nef59, Rab59g, Rit67, Sil65, Ste69a, Ste65b, Wal64, Ano96j, Ano97j, Ein59, Kor77, LS20, Rab56f]. Arthur [Ano91q, BGR61, Dan57, Fin65, For80, Kri85, Lam69, Mes83b, Mye87a, Rom82, Sto59a, Sto59b, Wol62, Woo65, Wri67, dF79b, Lan62]. Artic [Ano78n]. Article [Ano52q, Bel02, BF05a, Buk03, Cho03, Gre05, Hit03, Hol05b, Hol05a, Lug05, Moo03a, Pai02, Rab05b, She05a, SD55, Ure48b, Zha05, Ano50n, Jan78, McC54c]. Articles [Alb57, GR63, Jon63a, Rot03b]. artifacts [AH92e]. Artificial [Sha18, Gei17, HD23, Mec19b, Sep23, Tor19, Mec23d, Dan72, Smi54b]. Artillery [Hog68]. Artistic [Hol59, Ano93-27]. Artists [Har05]. Arts [Ano72f, Ano85a, Mor59c, Opp55, Why62]. Artsimovich [Fel73b]. Artyom [Zag91a]. Asa [Mil85c]. ASAT [Gar84a]. ASATs [Mon89]. ASC [Ano46d, Ano46-80]. Ashamed [Pip55]. Ashby [dF79f, And68, Rab48c]. Ashes [Rob62]. Ashley [Has57]. Ashton [For87]. Asia [Ano10a, Akh17, Arn96b, Aro51, AH22, Cal63, Car17, Cho86a, Cum01b, GN10, Gom75a, Jac80b, KL67, Kim22, Kla97a, Kla97b, KR12, Lal65e, Myr69, Pag68, Per78, Raj14, RR68, Sak84, Sar17, Sch58a, Sid00, SS88, Van22, Wei78c, Wei78d, Wio83, Zai06, Das96, Iri79, Lal67a, Sol68, Jac80b]. Asian [DeB81a, Ano82i, Ima93a, Jac55, Kli90, Mah59a, Per95, Dre82b]. Ask [AH89a, Rou75, Ano93r, McK25b, Moo15, Rot98d, NH08b]. asked [Bel20b, Rot94i]. Asks [Ano50-45, Ano52c, Mil62]. Aspects [Bet65, Bon50, Fel74b, Hug56, Ing60a, Lib59b, Lof51, Mac65a, Mar46a, Sci46, Sch47, Spe49, Ter58, Tin67, Ano85s, Fra76b, Tan84, Ada52]. Aspen [Gol70b]. Aspin [Ano92o, Isa85f, Isa87f, Isa92b]. Aspirant [GS09]. Aspiring [Ach12]. Assailed [Ano46-107]. Assassin [Got69]. Assassination [Har92a, SK69, HM15, Tob12]. assault [Gus11]. Assembling [Man74]. Assembly [Jac63a, May47b, AH91j, Ano47g, Ano50q, Fry57b, Jac67, Kih47f, Rab48f]. assess [Pug76]. assessed [BPG89]. assesses [Bet86]. Assessing [AM86a, Bla11, Ger20, GS09, Joh10, Lym11, SD10a, SD17]. Assessment [Ano71b, Ano84k, Ano90p, Dje72b, DWH83, Moo07, PF75, Sai72, Scl79, Cas78b, Jac79, PO16, Vog13, PF75, Fre84a, Gan83b]. Assessments [Mar65d, Str06g]. assessors [Hol83a]. Asset [Utl51, Wal21a]. assets [Bla11]. Assimilated [Rob75b]. Assistance [And54, Ano53-45, Ano54-31, Dan54b, Kin54, Lev72a, Mal54, Rev68, Woh54, All56l, Ano87-42, Lon82, Wyl56]. Assistant [Gol78a, Hor77]. assisted [You76a]. Assize [RMW54]. Associated [Ric74]. Associates [Pom67]. Association [Ano45n, Ano46-33, Ano46-30, Ano57a, Ano84k, Com54, Gla76d, O’B48, Sea56, Ano46e, Ano48f, Bea47b, Hod59, Kin80a, Lof47, Mic51, Pei50, Rad53, Wal55]. Assume [Ure48b]. Assumption [DuB48, Pri76b]. Assumptions [Nel00, Dud76]. Assurance [Gre87]. assured [RT12]. Assuring [Tur53]. Astin [Ano53x, Ano53-47]. Astrakhan [Ano64m]. astray [KGS25]. Astro [Haw73]. Astro-Archaeology [Haw73]. Astronaut [Sal71]. Astronautical [vBO62]. Astronautics [And58, Eds58]. Astronauts [Lew71b]. Astronomer [Bro20]. Astronomers [Mar61a, Ano95l]. Astronomy [Haw73, Lan56b, Lov63, Sak69]. asymmetry [Ano93-33]. ATBM [Mor87a]. Atheer [AH92k]. Athletes [Goo85a]. Atkin [Dro23m]. Atlantic [Mul65, RK93, Sig85b, You66b, Alb94c, Bar84, BW94, Bre69b, Mar64c, Mit83]. Atlantis [dF79h]. atlas [Hog97, Fel83e]. Atlases [FP90c]. Atlee [Ano45q]. Atmosphere [Bal79c, Dou60, Hel70, Jac73g]. Atmospheric [Ano62-27, Bro74, Fel62a, Fel62b, Kal62a, Rub62, Dro23e, LG15, Sto19d]. Atoll [Ano46-64]. Atom [AS00, AH93b, Ano48-30, Ano53-46, Ano54v, Ano56g, Bar64b, Bet70, Bor57a, Bor63, BF04, Bro48, Cha64, Dea52, Don86b, Ekl70, Gre55a, Hun85a, Ima70, Kap74, Kee56, Ken60a, Ken65, Kih47f, Lav55, LI70, Lil49b, Lil63, Mar49g, Opp51a, Rab67a, Ric74, Sag50, Sch80c, Sch54c, Sea71, Sim54e, Smi54c, Sto58, Ure56, Wal83b, Wil52, AJSB79, Bon57, Bon58a, Bon58b, Nat84, Ing70, Lan90d, Smi90c, Abb93, Bro47c, Fer57b, Ing70, Pat83, All56k, Nor04]. Atombombe [Thi46, Wei48]. atome [Bon57, Bon58a, Bon58b]. Atomic [Ato50, Abr73, AM86a, Ack90b, Ada52, Aeb48, All54a, All55a, All57k, And56, And65, And71, And10, Ano45a, Ano45e, Ano45f, Ano45g, Ano45h, Ano45i, Ano45o, Ano45p, Ano46-98, Ano46f, Ano46i, Ano46h, Ano46k, Ano46g, Ano46j, Ano46-100, Ano46m, Ano46-104, Ano46-101, Ano46-102, Ano46-103, Ano46d, Ano46p, Ano46w, Ano46v, Ano46t, Ano46y, Ano46z, Ano46-45, Ano46-32, Ano46-39, Ano46-42, Ano46-43, Ano46-106, Ano46-46, Ano46-49, Ano46-50, Ano46-60, Ano46-67, Ano46-71, Ano46-87, Ano46-93, Ano46-95, Ano46-92, Ano46-90, Ano46-94, Ano46-107, Ano46-105, Ano47c, Ano47b, Ano47d, Ano47f, Ano47l, Ano47n, Ano47r, Ano47t, Ano47u, Ano47w, Ano47y, Ano47-27, Ano47z, Ano47-28, Ano48o, Ano48g, Ano48j, Ano48k, Ano48l, Ano48m, Ano48n, Ano48h, Ano48f, Ano48e, Ano48-28, Ano48-29, Ano48-31]. Atomic [Ano48-35, Ano48-39, Ano48-41, Ano49k, Ano49j, Ano49r, Ano49-27, Ano49-29, Ano49-30, Ano49-39, Ano49-40, Ano50g, Ano50l, Ano50v, Ano50-35, Ano51b, Ano51l, Ano51d, Ano51r, Ano51y, Ano51-28, Ano52b, Ano52e, Ano52t, Ano52-27, Ano53a, Ano53b, Ano53c, Ano53d, Ano53j, Ano53k, Ano53-43, Ano54a, Ano54f, Ano54o, Ano54q, Ano54t, Ano54u, Ano54-30, Ano55d, Ano55l, Ano56e, Ano56j, Ano57e, Ano57c, Ano57b, Ano57k, Ano58b, Ano58l, Ano59b, Ano60g, Ano60f, Ano61e, Ano62m, Ano64h, Ano65d, Ano66a, Ano66f, Ano67c, Ano68b, Ano68l, Ano69b, Ano69d, Ano71a, Ano72d, Ano73a, Ano74a, Ano75b, Ano75g, Ano76a, Ano77k, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano78h, Ano78a, Ano78b, Ano79b, Ano80e, Ano81d, Ano82b]. Atomic [Ano83e, Ano84c, Ano85e, Ano86m, Ano87t, Ano88f, Ano89h, Ano90q, Ano91t, Ano92k, Ano93o, Ano94i, Ano95h, Ano96h, Ano97h, Ano98i, Ano14a, Arm53, Arn47, Aro50, Axe51k, Bac48, Bar58a, Bea47b, Bec58, BAB^{+}46, BH62, Ber93b, Ber03, Bet54, Bet62a, Bet77b, Bha55, Bla46, BBD^{+}46, Bon51, Bri87, Bro60a, Bro47a, Bro47b, Bro49, Bro47c, Bru50, Bru57, Bur50, BKSG60, BBHGA57, Cou52b, Cah76, Cap98, Car49b, Car49c, Car49d, Car49e, Car50a, Car50b, Car50c, Car50d, Car50e, Car50f, Cas77e, Cav47, Cav50, Cav52, Cav53, Coa47, Coc55, Coc59, Coc51, Cog82, Coh53a, Col59, Com57a, Com46, Com53, Cus48, DB52a, DB52b, DB52c, DS47b, DB50, Dan46, Dav51, Day74a, DeN78, Dea51, Ded60, DR71, Deu60, Don89, Dow82h, Dow83e, Ebe47, Eli81, Eme60]. Atomic [Eme63, Fed59b, Fel55, Fel72a, Fel83a, Fer57a, Fin54a, Fin54b, Fin55, Fin57c, Fis46, For86, Fox46, Fox48, Foz61a, Foz62b, Fra61b, Fri48a, Fry56a, Gav51, Gla56b, Gol62a, Gol62b, Got50, Got51, Gre62, Gre67b, GR63, Gro47c, Gro49, Gué70, Gus83, Haf49, Hal07, Ham61b, Ham61c, Han46, Has57, Heb52, Hei47, Hei68, Hei70, Hin53, Hir50, Hod55, Hod59, Hoe88, Hos87a, Hos87b, Int50, Ing52a, Ing52b, Ing53, IW49, IW52, Joh81a, JCKG^{+}50, JC50, Jol46, Jon59, Jos61, KW15, Kal83b, Kay49, Ken62, Kih47a, Kih47b, Kih47c, Kih47d, Kih47f, Kih47e, Kih48a, Kih48b, Kih48c, Kih48d, Kih48e, Kih48f, Kih48g, Kih48h, Kih48i, Kih48j, Kih49, KFM50, Kno55a, Kno55b, Kow48, Kra59, Lan56a, Lan56b, Lap48, LB53, Lap54b, Lap54c, Lap57, Lap59a, Lau47, LeR50b, Leg56, Lei47]. Atomic [Ler49a, Lev51b, Lew50, Lie46, Lil46, Lil47a, Lil47b, Lof47, Lof51, Lun59, Mal63, Man95, Mar59a, Mar49a, Mar83c, Mar00, Mar46a, Mar46b, MNS46, Mar48c, McC96, McC48, McC56, McM49b, McM50, McM51, McM54, McM67, Mec20e, Mei48, Mei56, Mes87, Mic55, Mic51, Mil47, Min48, Mor56, Mot48, Mot49, Mur54c, Nel75, New54, Nie63, Noy46a, Ols62b, Oom46a, Oom46b, Opp46, Opp48a, Opp53, Opp63a, Opp63c, Osb47, Osb49a, Osb49b, Pal00, Par46a, Par66b, Par71, Pei50, Per55, Phi54, Pic51, Pot49, Pou78, Pri56, Pry48, Pya74, Qua80, Rab47n, Rab47a, Rab47h, Rab47b, Rab47d, Rab47k, Rab48f, Rab48e, RG49, Rab50b, Rab51a, RS52, RDS^{+}52, Rab56e, Rab62a, Rab68g, Ran50, Ran53, Ran59, Rei50, RKR55, RK55, Rho46, Rus46]. Atomic [S.46, Sch47, Sch49, Sch65b, Sch69d, Sch98, Sei49, She60, Shi46a, Shi47b, Shi47c, Sho46, Sim54d, Sim54e, Ski53, Smi54c, Smi58d, Smi68, Smy50b, Spi66c, Spi50, Spo50, Spo51, Sta48, Ste75a, Ste52, SL46, Sti47, Sti48, Sto58, Str58, Szi47a, STS^{+}49, Szi50a, Szi60a, Tel47a, Tel47c, Tel49, Tho53, Top58a, Tur51, Tur52, Upt63, Ure46b, Ure47, Ure48a, Ure49a, Utl51, Van46, Van52, Wal90b, Wea82, Wei48, Wel08, Wig51, Wil53d, Woo92, Wri46, LLR^{+}47, You76a, ZA04, Zlo47, dR49, vN57, vW57, AC02, Ack89g, Alp85, Ano47k, Ano86o, Ano98m, Ano15f, Aue03a, Cam88, Che52, Ebe20, Gow79, HRS81, Hun04, MSW20, McK83, NB77, Rab50e, Rot77, Rot92n, Sha87c, Ste77c, Thi46, All53, Ano45b, Ano46j, Ano46l, Ano46-65, Ano46-66, Ano47e, Ano47j]. Atomic [Ano48i, Ano48w, Ano48-34, Ano50c, Ano54c, Ano54u, Ano73d, BW59, Bey46, Bru47, Can63, CH54a, Dew48, Haf51, Haf54, Hig66, Hus53, Kal54, Kal63b, Kuy54, LeR49b, Lev46, Lev47, Lev48, Mar48b, MT55, McL53, McM48a, Mil48b, Mil52, Oom46b, Pei70, Rab49a, Rab49c, Rab52a, Rab59j, Rab60a, Rab73f, Ran55a, Sem59, Sim54e, Spe49, Tel48b, War48, Wea85, Ano50-41, Boy05, Bro57a, Bro57b, Cav57, Dan57, Fra61a, Gre71, Kat53a, Kat57a, Lap55a, Lap57, LeR50a, Moo00a, S^{+}98, Sha50, Soc54, Sti52].

Atomic-Powered [Mar00]. Atomica [RK89a]. Atomized [Rab59g]. Atomphysik [Thi46]. Atoms [Ano55t, Ano78m, Aue03b, Bro47c, Che82, Gal76, Gar93, Gre82a, Lop78, Mar55a, Rab48b, S.55a, SR80, Wei03, Zei64, Fot89, HD23, Kim22, Lev12, Smi58f, MM82, Mar82c, Bro47c, Cam54, Fin57b, Hug57, dK60]. Atoms-for-Peace [Ano55t, Zei64]. Attachés [For57]. Attack [Ato50, Ano48o, Ano50-41, Ano62p, Bro60b, Cav53, Coa47, Lal64g, Sch59a, Ack88c, AB18, Arn22, DE23, HBD17, NA88o, Ruf90, Sig10, SK85]. attacking [Ano88w]. Attacks [Mar62, vH85, Bre13, Sme22]. Attali [Hib95a]. Attempt [Ram88b, Sha90]. attempts [Yor75b]. Attention [Ano91v]. attic [Ano98p]. Attitude [Khr05, Sie52, Tay52, NA90l]. Attitudes [FD76, Lal67f, Lew71e, Bin18, PD12, Suz84]. Attlee [Atl47]. Attracting [Ger13, Ibr13, Zer13]. audacious [Hen17]. Audience [Mas08]. Audiovisual [H.84b]. audit [Sch98, S^{+}98]. Auerbach [Lan56a]. Auger [Ano46-78, Gal54, Nie55, Rot62, SD55]. augments [Bur88]. August [Ano55l, Bev56, Dow80h, Dow82h, Dow83b, Ein48a, Jac92, Ler80a, Pop52, Fat46, Ano80b, Ano80c, Ano80d, Ano81b, Ano86v, NA98e, Ram82, Top56]. Augustin [Ada52]. Augustus [Nei82a]. Aura [Bal79c]. Austen [Kih47f]. Austin [Tra71]. Australia [Ano48n, Bro74, Cam79, Mil87, OW15, Sut04]. Australian [Len14]. Austria [Ano58l, Ano58k, NW79, Pat79b]. Austrians [Gom75c]. Authentic [Jef94]. Author [Alp92, Ano97g, Bal60b, Fel66a, HY67, Mar82b, McM48a, Per97, Ano98n, Dro17, Dro23j, Dro23k]. Authoritarianism [S.80, Tos82]. authorities [BV20]. Authority [Jas81]. Authors [Ano61a, Ano62a, Ano62b, Ano62c, Ano62d, Ano62e, Ano62f, Ano62g, Ano62h, Ano62i, Ano62j, Ano63a, Ano63b, Ano63c, Ano63d, Ano63e, Ano63f, Ano63g, Ano63h, Ano63i, Ano63j, Ano64b, Ano64c, Ano64d, Ano64e, Ano64f, Ano64g, Ano65a, Ano65b, Ano65c, DT91]. Auto [Cic74, Mus79, Sch70a]. Autobiographical [For80, Hah67a, Hah67b]. Autobiography [Her90a, Kil78, Sto68a, Inf06, Wil77b]. Automated [Gei17]. Automatic [Lan56b, Mei54, Rau69, Ste69b]. Automation [Bat69, Deu68, HSJH18]. Automobile [Cor69, Wou81, Wou73]. Automobiles [BF73, Bro84a, Wou81]. Autonomous [Zim75, Cha16, HCH16, Rof16, San16, Aue78]. Autonomy [Fis83, Ark87a]. Autumn [San66b]. Availability [Ano46n]. Available [Ano49h, Han98a]. Avenger [Isa91e]. Avengers [Rab60d]. Avenue [Gou75b, Isa90m]. Average [Ano87j]. Averted [Joh00a, AG98, DP12]. Averting [Cho86a, Per87b]. Aviation [Hin69b, How63, Kir90, Rau69]. Avital [Dow80i]. AVLIS [Ano88l]. Avner [Par86, Moo94b]. avoid [Ano14g, GvH23, Leb85, vH^{+}90b]. Avoided [Bus47, Kin84b]. Avoiding [Ram05, SW21, Wal85a, Wei82a, Bla84]. avoids [Cai88]. Avorn [Lyn69]. Await [Hel02]. awaits [Nas18]. Awakening [Ber72d, DP20]. Award [Ano68g, Ano83o, Ano85h, Ano86n, Ano87z, Ano91u, Ano51o, Ano52-28, Ano65f, Ano85u, Ano93-28]. awarded [Ano92-32]. Awards [Rei63]. awash [GM14]. Away [Cle04a, Gal90a, HM90a, Mue01, Nie66a, She74a, Ano92r, Ano94o, Bar80l, Gal90b, Gur88a, NA88l, NA90n, NA90t, Wit82, Zel76a]. Awesome [Bai75a, Fol74, Bai20]. Awestruck [Arn91]. AWOL [Ard93a, Isa03c, Isa03g]. Axing [NA90p]. Azbel [Ses83]. Azerbaijan [Yun94].

B [Abr74, All54a, Ano46-70, Ano78l, Ano78o, Bal79c, Cal82, Cal84a, Cle95, DeN74a, Dob54, Fol74, For87, Gil85, Gol88a, Gol95a, Haw47, Hob85, Hur78a, Jac80b, Kat54e, Key81, La 91, Lep89, Mar54e, Nef78, Nel84, PF75, Rat83, Sag86, Sar85, Sch80c, Sha59, Sim87a, Sti83, Tir88, Tsi72a, Wil71, Ano46-52, Ano93-31, Ano99c, Bez84, Bro90, CP93a, Day76e, Fed68a, Hog86, Hyd91, Isa89i, Isa90u, Isa91j, Lan90m, NA89i, NA92a, NAB99a, Pie19, Pik96, Rab62a, Ram82, Sch76a, Ste89d]. B-1 [Bez84, Day76e, Hog86, Sch76a]. B-2 [Bro90, Isa89i, Isa90u, Isa91j, Lan90m, NA89i]. B-29 [Ram82]. B-2s [Pik96]. B-52s [NA92a]. B-70 [Rab62a]. B61 [KN14a, NKH03a]. B90 [DH90]. Baade [Asi62a]. Babai [Hur78a]. babble [Mar98c, Sho92c]. Babies [HG71, Rus70a, BV20]. baby [AM21, GM90]. baby-boomer [GM90]. Babylon [Opp62a]. Baccanalia [Rau91]. Back [Aft00, Ano46a, Ano97n, Bur74, Cam60, Cho95, Cle04b, Eba72, Fry57b, Hub60, Isa99d, Isa01b, Lan90g, Mei64, Myr75, NA89c, Rab73b, Ric01b, Ric03a, SSK10, Tel50, WD03, Woo04, Ano90-61, Ano96n, Dav55, Dro23i, Fel76i, Fel79d, LKWR89, Lyt85, Man15, MM98, Mat84, Mec21h, Mia98, Obu89, PGL91, SM82, Wei85c, dW90a, vH77b]. Back-Door [Myr75]. Backdoor [Gho15b]. Backed [Jac74b]. Background [Bla82a, Fro49, Mes70, Sak69, Bas86, Mat98c]. Backing [Bul09a, Wit82]. Backlash [Jac74d]. backpacks [Ark85d]. Backs [Ano46-45, Tru49, Ano93u, Zac92]. Backward [CLR75, Mar54d]. backwards [Fel81k]. Backyard [Sel69b, Tuc01, Ano91-40]. Backyards [Dow82e]. Bacterial [Ame47]. Bad [Ano91w, Ark94a, Cas04, Cum01a, Fly00a, Gol90d, Ist98, Kev07, Kra15, Pra05b, Rob08, Rot03b, Sto78b, Weg96a, Zam96a, Ano90-47, Bar83b, Fel77d, Gla85, Har89b, Isa89m, Isa97a, Kis79a, Koe22, LP97, NA89i, Rot96e, Sch20, Spe85b, Woo97b]. Badash [Fel81a]. Baden [Ano59d, Ano10a]. Baden-Baden [Ano10a]. Baeyer [Ano94h]. Bagdikian [Goo84]. baggage [Pow90c]. Baghdad [Car81, Gom81b, Pra05a]. Bail [Com74]. Bailey [Bou98b]. Bailing [Lew72d]. Bain [Bre74]. Bait [Ano91-41, Ano91p, Ano95i, Moo05b]. Bait-and-Switch [Moo05b]. Baiting [Con58, Leb91a]. Bajpai [Das96]. Baked [BB63]. Baker [Deu83b, Moo94i, Nor64, Wei94b]. Baku [Lon98]. Balance [Bin67, CF74, EB73a, Fel67a, Jan76a, Lon68, Ols74a, Rab72d, Sei62, She07, Voy76, WH00, Cum97a, Dav76b, Gar77b, Kap93, MH89a, Yor76a, Eng88]. Balance-Wheel [Jan76a]. Balanced [De 74, Pit64, Los54]. Balancing [Atr06, Ark83c, Ark87d, Woo86, Ano88e]. Balderson [Ano50-41]. Baldwin [Ano46-44, Bal50]. Balkan [And85]. Balkans [RP82, Woo96c]. balks [NA88m]. Ball [Dow80f, Fel76b, Men86, Moo98e, Rot99a, Ano46-93]. Ballistic [Cof65a, Dys64, Jac56, Mar67b, McM65, NK09c, Stu89, Cot19, Kar88c, Kar88d, KK19a, LvH18, Rub90, Dys72]. balloon [Gou23]. Balloons [She83a]. ballot [Bar77a]. Baltic [Chu94a, Han90a]. Ban [AM21, Ano50t, Ano56i, Ano56m, Ano60r, Ano62-28, Ano63n, Ano63r, Ano63s, Ano63t, Bec64, CP99, Con60, Dav60, Dys59, Eps88b, Fed58, Fel88, Foz61k, Foz62i, GSB^{+}83, Gla02, Gue14, Hev47, Hla65, Hum59a, Hun90, Ing54a, Ing59a, Ing60b, Ing63b, Ing64, Ing79b, Isa99f, Isa03f, Jeh60a, Kal95b, Kar14, Kas21, KK63, Kib17, Kim01, Kol64, Lal66g, Lal67b, Len14, Mar83e, McN63a, Nel09, Ore59, Ore60b, Pif24, Pre63, Pre17, Rab57d, Rab59m, Rab63j, Ros59, San63, Sin17, Spi59, Sto97, Tur60, Ure61, Ven99a, Wes58, Wes85, Wil83, You61, Zam93, Zam96c, Zer13, Ano86-30, Ark86h, AS94, Bro88b, Bur90b, Cha16, Cle87b, Cle89, Col93b, DM81b, DM85, DM86, DT91, DvH91, Eps86, Eps90a, Eps91, Eve85, Eve88a, Eve88b, EG89, Fet87, Gaf86, Gol88f]. ban [GH83, Gub15, Isa90s, Isa98e, Isa98f, Jeh60b, Joh95, Joh96b, Kid91, Mat83a, McK86, Mec20a, Mia16, Moo97l, Moo97p, Nag15, Pod24, Pom77, Pom79, PE89, Rof16, Ros87a, Sal87, San16, Sau17, SL87, Sip86, Tov89, Ven99b, Wit82, Wur96c, Wur96b, Yor81, Zam92d, ZK94, Zam98, Zha15c, vH91, Ano79e, Wie82b, Ano90n, Lon82]. Ban-Wagon [Isa03f]. Band [Fin57c]. bandits [Ark85b]. Banesh [For80]. Banff [Pug81]. Bang [Ano74e, Ano97m, Col93b, Gor96a, NA88k, NA89j, Rau92, Van90b, Ste74]. Bangladesh [Huq11, Kap72]. Bangladeshi [Huq16]. Bangor [Dow82d]. Bangs [Sch75c]. Bank [Fay75, Ano96m, Bou99, Coh69]. banken [Joh96a]. Banking [Woo00f, Ark87g]. Bankrolling [Dis04, Kel03]. Banks [LK10, Usm79, Zag94b]. Banning [Eps79, Kin78a, Kib17, Mit58, Pre17, Sin17, Wes79a]. Bans [Ano63u, Mti96a]. Baptism [Ste75a]. Bar [Moo02a]. Barbara [Mei77a, Nas62, Sto55b, Ven80]. Barbarism [Sma76b]. Barden [Bar64a]. Bargain [Fry56c, Goo77, Sin05b, Asp83, Har76a, Loe86, Sch84a, Tro79, Tsi83b, You84a, Ano96n]. bargaining [GB77, Lan10, NA87h]. bargains [LN83]. Barghoorn [Ham61a]. Barker [Fer69a]. Barnet [Mul65, Wil84, Gol69b, Mar79b, Ros90a]. Barnett [Iri79]. Barnouw [Dow80h]. Barnum [Sto76b]. Baron [Ano87r]. Barons [Her82b]. barred [Col88]. barrel [Har86d]. Barrels [BC16]. Barrett [Rus51a]. Barricades [Pla69]. barrier [Fas78, Kid91]. Barriers [Gif13, Sei71a]. Barry [Gri83a, Hob85, Mil86b, Mor82a, RK89a, Roc81, Smi83, War86, Ano86v]. Barth [Kal62b]. Bartimus [Mos92]. Bartlett [Sin81]. Bartocha [PF75]. Barton [Alv87, Iri86, Mar78a, Ano50-36, Ber87b]. Baruch [Ano46p, Ano46o, Ano46-79, Dre46, Lac90, Pes87, Sme74]. Barzun [Lyn69]. Base [Smi75a, Ano95i, Isa93h, NA93i, NA94f]. base-closure [Isa93h]. Based [BAB^{+}46, Ros75, Tay87, Ano18, Gre18, Rob18, Ste87, Wol17, Woo12]. basement [Ano92-28]. Bases [Kea60]. bashing [Lan90m, Rot93m]. Basic [AL62, Ber72a, Bru58, Cre74, Cro58, Joh65b, Kay66, Kis66, McK74, Mor67, Neu58a, Pei53, Pom65, Rab66a, Sub70, SL58, Smi93a, Wat57, You58, Lap59a, Pla60]. Basics [Sma77, NA89c]. Basis [Ehr81, Gro51, Roc18]. basket [NA91i]. baskets [Zim81a]. Batavia [Gol69c]. Batchelder [Fer62]. Bateman [Ras55]. bathtub [Low93]. Batman [Rot93k]. bats [Ano94k]. Batteries [MS06]. Battery [Wou71, Dro23a, KK18a]. batting [Ano97j].

Battle [Mor82a, Mul53, Bul02b, Wal85b, BG14, Dav92, Esp14, GMV^{+}24, Hun14, Isa90a, Man95, NK12a, Thr92, Tom14, Zag95c]. Battlefield [Ano79a, Raj14, San14, Dow81d, Dow82h]. battlefields [MK02a, Van03]. battleground [Mec22d, Dow80h, Bou98b]. battleships [Isa89l]. Battling [VSD^{+}09]. Batty [Bar80m]. Batzel [Bat88]. Baubles [Ano91x]. Bavaria [Gla02]. Bay [Isa89l]. Bazaar [Har91b, Kal78, Kup06, IG91, Kla85, Rot97h]. BBC [You68b]. Be [Ano92s, Ber49, Bro72, Bus47, Dje74, Eis66, Fed59a, Fel67c, Fri48a, Gou75b, Gué67, Kah59, Ken60b, Kin67, Kou75, Lev47, Lev48, Lil46, Lym08, Mel75, Mel58, Mil85b, Pan00, Pol49, Rab62h, Rob75b, Rob08, Rot61, Rot04f, Sie05a, Spi59, SK03, You69, All17, AA54, Ano46i, Ano46u, Ano46-82, Ano95j, Ano97i, Ano14g, Ano15e, Ark83a, Bar77b, Bar81e, Bel20b, Bet46, Bet20, CW23, Cla92, Cla20, Cla77, Cli77, Coa88, Col79c, Dah77, Dav55, Eme61, Eps94, GLH^{+}79, Gre12, Har89a, HS13, Hay79, Hec12, Hec20, Her84, Ing76, Kal81, Kel90, Kin84b, Kor19, Kre88b, LP10, MSW20, Mec20b, Mec20h, Mil79, Moo25, Moo15, Mor47b, Mos67, NH08b, Nov77b, Pra22, Rab51k, Rao87, Rev21, Rho15, Roc76, Rot77, Sch21, Sez20, Sho80, Ste77c, Tay21, Wol80, Zal19, vH79]. Beach [Pug02, Mac81c, Nor82, Doh58]. Beaches [Woo00f]. Beagle [Lew72a]. beam [Gar78b, GR79]. bean [Ano87-31]. Bear [Kus83, Ano95n]. Bearing [Dje74]. Bears [NA88b, NA90d, NA90n]. Beast [Kis96, Kli59, Hul76]. beat [Ark97f, Lyn78]. Beating [Alb94a, NA90a, Wur97]. Beaton [You66a]. Beatrice [Mec20a]. Beaufre [How65a, Mas66]. Beauport [Ano58g]. Beautiful [Ano52z, Edm74, Tor77c, Wil73c]. Became [Bor65a, Wea11]. Bechhoefer [Cav61]. Becker [Rei72]. Become [Mor56, Kib12, Mec22d, Nay12, Squ12]. Becomes [Day75h, GB11]. Becoming [Met24]. Bed [Cha88b]. bedfellows [Isa98a]. Bedford [Abb91]. Bedlam [Wor63]. Bee [Szi65]. Beef [Jon93b]. Beeler [Ano60f]. Been [Che52, Com46, Gla75a, Pri76a, Lon76b]. Beer [Ano94j]. bees [Ano97i]. Before [Day77b, HPL^{+}49, IF51, Mot47, Rab46b, Rab47g, Rab58e, Sap75b, Ano86v, Bev56, Che52, Chu54b, Eis54c, Kal54, Kal63b, Ksa91, Mec20d, vN57]. began [Wei17]. begets [Ack87d]. Begin [Cho81, Sha97, Ano94m, Fel77j, For91, Mia16]. Beginners [Dow84d]. Beginning [Bro47d, CF46, Mey66, Rab63g, Sin67b, Sto67, Wes00b, Gil18, Kal81, Oli82, Sco08b, Smi92a, WSG^{+}15]. Beginnings [Wil77b, Fel78l]. Begins [Opp02a, Ada83, Isa78, Smi93b]. Begun [Rab73d]. behalf [BL81, Rep83, Woh78]. Behavior [Mic55, Nor64, Soy89, Tho80g, Wyl56, Gri84, Pri70]. Behavioral [Mea69, Nor64, Nor64]. Behind [Ano56d, Ano86f, BJCP88, Brz89b, Fel75b, Han92b, Hus03, Llu00a, Mes70, Moo96a, NCA88b, Smi58d, Str06a, Tsi72b, Car03, Dro23l, Dro23m, LS94, Llu00b, LC88, Mil10, Klo85]. Behrens [LeR50a]. Beijing [BV20, Zha05]. Being [All50, Ano91-33, LL90, Sch80d, Woo97a]. Beings [Arn54b, Fre72b]. BEIR [Bio73, Com73, Kat73, Tam73a]. Bela [Low93]. Belarus [Mar93b, Tiu94].

Belfry [Sim54e]. Belgium [Zan91]. Belgrade [Jac61]. Belief [Ano50c, Bor57b, Hoa77]. Beliefs [Kol75, Fie24]. Believe [Ano47x, PMRP50, Vis56, ZS15]. Believers [Ano98o]. believing [Bra13a]. Bell [Ros67]. Bellicose [Sch06a]. Bello [Pen52]. Belongs [Sho92b, CO14]. Below [Lla70, Ano98l]. Belt [NA90l, Sah20]. Beltway [Ark85b, SM89a]. Ben [Dow81g, Goo84, Pee78]. Ben-David [Pee78]. bench [Sch89a]. Bended [Wei89a]. Bender [Ano81f]. Béné [Ano60f]. Beneath [Chu03, Wil76b]. Benefit [Ano78o, Hol83a]. Benefits [Ehr75, Mac88, Bug22, DS47b, Flo78, Lin85]. benevolent [Sch84b]. Bengal [Rab72d]. Benign [Fre02, Her95b]. Benjamin [Ano87s, Lon82, Mei90, Sch70b, Sto59b, Aur71d, Dow82c]. Bennett [Coo81]. Benson [Dys72]. Bent [Pat81a, Wel60, Ste89f]. Berelson [Nor64]. Beres [Don86a, Gan84]. Beria [Rot93v]. Berkeley [Ney73, Alv74, Goo55, Lyn69, Mar61a, Sch69c, Sch78b]. Berkhout [Rot93a]. Berkner [Adv50]. Berlin [Bou98b, For80, Khr63, Ano91-39, Fle59, Har61, Lan61b, Nea61, Rab59a, Rab59c, Rab61b, Sti61, Szi60b, Tuc64]. Bernal [Nor06, Sto59b, Bro05, Haw56]. Bernard [Bus85, Coh69, Gab73, Gol83c, Ing82, Nor64, She60, Win81a, Ano75a, Ano93p, Rab73d]. Bernardo [Fly00a]. Bernhard [Cav61]. Bernie [WRT93]. Bernstein [Ber87b, Mor79, Sno68a]. Bertin [All56k]. Bertrand [Fel83a, Ing56c, Kap61, Rit67, Sto68a, Ano80o, LB68]. Beryllium [Bel02]. Beschloss [LaF86, Leb91a]. Beset [Cla74]. Besieged [Wal83b]. Best [Ano75f, Ano91-34, Che89, Gil83, Kos04b, Spo51, Cot19, Isa89m, Jer85, LSP95b, Man15]. Bester [Ten75]. Bestrides [Rab67a]. Bet [Bet77a, Dro24d]. beta [Moo95a]. Bethe [Dow83a, All59a, Ano85u, Ano92-32, Bet93, McD62]. Betsy [Jan59]. Better [Ano60q, Ano96l, Ano97l, Arb91b, Aue06a, Bet89b, Cir00, Con60, Eps77a, HvH75, Kor75, Lug00, Pea90, Ros75, Bul08c, Ste75b, Wei68c, AB77, Ano96o, BP95, Cas78a, Cla90, Cou81, Cri94, Dro24e, Gar97b, Gol11, Kis76a, Kre95, Lan89g, Mar84, Hir88]. Betts [Eng88, dF79d]. Betty [Ano88e]. Between [Ano63r, Ano95m, CF74, Dil66, Eké21, Nor64, Pei69, Por71, Rab71a, Rep03, Sag92, Woo56, Ano51l, Bol88b, Cah94, Cle11, Dub55a, Dub56a, GS96, Hoe87, Mec19f, PR14, Pod16, Sig10, Bla84, Bre93, You82]. Beuret [Pat81a]. Beveridge [Ing56c]. Beware [DMMH02, Gof93a, Han01]. Beyerchen [Gou78b]. Beyond [Ach10, Ano99d, Cum04, Dre77, Fot91a, Gre68b, Ham91c, Hei70, Hol74, Jac58, Kor90, KH13, Lan76, Mei56, Moo93d, Ner80, NA92e, Pea91, Rab61b, Spa69, Swe89a, Tho80d, TY75, Tur13, Tur14, Ano14e, Das96, Hei71, Sco08b, Sim90b, Van20, VU22, Bri86, Gus83, Hol91, Abb84a, Bla84, Der88, Fre76a, Sol68, Tuc72]. Bhabha [Per00]. Bhalchandra [Moo93b]. Bhupendra [Ano79a, Deu83b, Lac85]. Bialer [Fis83]. Biased [Zei73]. biases [Kap95]. Bible [Koh76]. Bibliographic [Ano87q]. Bibliography [Ada52, Ano48p, Ano50-44, Ano62p, Lei47]. Bid [McM50]. Bidding [Ano92y, Loc92a]. Biden [Fie21a, Gol21, HK21, Kas21, Kor21, Meh21, Rev21, Squ21b, Vad21, Zik21].

Bienen [Dev69]. Big [AH01, Ano74e, Ano98n, Arn98a, Fre68b, Fre77a, Gou61, KAT^{+}71, Liv69, Lor02, Mot68, NA90m, Pal01, Pet78, Rot92c, Sch75c, Tel71, Van90b, Wal90e, Whe84a, Aba89, Bar83b, Dro21c, Dro21d, Dro24f, Gho15a, Isa94g, Klu22, Not15, Obe97b, Sch75e, Wie87, Wil15, Zam95a, Ano55q, Atl47, Joh94b, Kam74, Ste74, Ust93a]. Big-Bang [Ano74e, Ste74]. Bigeye [NK86b]. Bigger [Jac74b]. Biggest [Ler53, Hus24, Klu22]. Bikini [Ano46h, Ano46q, Ano46-64, DuB46c, Kis54, Lau46, Nor94b]. Bilateral [Ben61, Rev68]. Bill [Ano46r, Ano46-61, Ano48q, Ano55b, Ano65h, Day74c, Day97, Dow80f, Dow84c, NH08b, Rab52a, Sti83, Ano91y, Med91, Rob88, Str88, Ano45k, Ano46-55, Ano46-54, Ano46-56, Ano46-85, Ano46-91, Ano46a, Ano47o, Ano50-43, Ano12a, Kib04, Lef77b, McK23a, Pri47, Cou50, Sha94, Tru46, Tru50, Wal46b]. Billion [Pai04f, Pra95, Rea80, Rea68, Ros75, Sha00, Spa95e, Wad74, Isa89h, Isa93a, Moo18, Rot96b, Ano96o]. Billion-Dollar [Ros75, Isa93a]. Billions [Har99, McP65, Arn22, Pos21, Ust93d]. Bills [May47a]. Bin [ABH02]. bind [Pow90a]. Binder [Ano78n]. Binding [Fol74]. Binge [Fel66b]. Bio [Ano99e]. biochar [Sto19a]. Biochemistry [Kat54d]. Biodefense [Sch05f, Bul08b, Wri04, Ger17, Kir17]. Biodevelopment [SD09]. biodiversity [Hen17]. bioenergy [SMRVS22]. Biography [Gol88b, Lan98, Sha63, Dro23j, LS94]. Biological [Ano50l, Ano52c, Ano82d, Ano87k, Ano91s, Ash60, Bar61b, BHH11, Bra89, Bru48, CT78, Chi59a, Chi59b, Dan72, Dje72a, E^{+}84, Fed64, For49, Fre95, Fre72b, Fri72, Gla60, Gol72a, Gro84b, Hoa72, Ims60, Kel98a, Kli81, Kre89c, Leg92a, Len11, McL11, Mer46a, Mes70, Rab59i, Rab59n, RH65, Rab72b, Ram72, Rob75b, Sei71b, Sel59, Sid69, Sim11, SD10a, Tuc97, VT73, Wil73b, Wol55, Ano68l, Ano87y, Ban11, BBP14, Ber87c, BWL21, Ehr84, Fra21, GPR81, HIKWB10, Len20, Rap11, Ros87a, Smi16, WS83, WSH^{+}84, Wri87, Wri89, Ano72b, CK20, Mes64a, Kli81]. biologischen [Ano68l]. Biologist [Bli70, Mul67, Pit67, Tho80g, Dro20b]. Biologists [Thi47, Wad63]. Biology [A.54, Ano69l, Bea64, Dow84b, Eis72, Gla62b, Kam46, Lla70, Lou51, Nor06, NL61, Pit61, Rab52e, Rab65h, Ros52, Sti60, Tho80g, Ano78p, Gou95, Hoa77, OGFM18, Tuc10b, Wri85, Ano78p, Olt52]. Biomass [Bis14, DR14, Mor14, Boo22, Koe22, Pie22]. Biomedical [Ber72a, Fel83d, Kap76, NBM85]. Biorklund [Ano57c]. Biosecurity [Ano07r, BST02, Eps12, Gro15, MZ07, SD09, Tuc10a, Ano12d, VA11]. Biosphere [Hal75a]. Biotech [Ano90s, Rot89b]. biotechnology [Vog13]. Bioterrorist [Dan05, Dan20]. Biowar [Wri91, Ano90s]. Bioweapons [Cho02, Fre95, KND10, Mei01, Nix10, Tuc08, WD03, Wri97, Wri02, Isa90g, JH16]. bipartisan [Gor84]. Bird [Dow83e, Ano94j, Ano97j, Rot92j, Dro23k, NA87d, Fel86a]. birds [Ano97i]. Birdwatcher [Lew72c]. Birth [Ano49g, Cal62, Cho58, Coh69, Com53, Ded58, Dje74, Dje81, Jac74b, Lev07, Sko67, vHJ97a, Dav86a, Kuz88, McM05]. Birthday [Opp79a, Ark87e, Har79c, Mec15b, Rot97g]. Birthplace [Kam74]. bis [For80]. bishops [Mil83a, Mil84]. bit [Ano88g]. bitcoin [McK22, Pra22]. bites [Hia87, Zag92]. Bitter [All59l, Edm74, Rab75, Ste88]. bitty [Ano97n]. bizarre [Rot97h]. bizniz [Ano94j]. Black [Lal67a, Moo04, Myr72, Ner69, NA88a, Pas76, Ric01b, Ano91-35, Ano92v, Ano96j, Arn92b, Bel94, Bin79, Kla88, Law11, Rot92j, Ade82, Ano63u, Ark88b, Coc89, Law11, Sim87a]. black-market [Bel94]. Blacker [Win84b]. Blackett [Ano49-30, Bro57b, Cav57, McM49a, Mor49, Pan63, Shi49]. Blacking [Ano95j]. Blackjack [NA92k]. Blacklisting [Wit95b]. Blackmail [Eng88, DSL^{+}87, Ger84, Gil98]. Blackwood [Dow83a]. Blair [AM86b, Lap54c, Pul04]. Blam [Ano97l]. Blame [Ano98o, Ark97b, Bou02, Zha05, Ard93c, Kar97]. Blames [Kih47f]. Blaming [Loe02]. Blanca [Jag98]. Blank [Isa04c, Ano92-34]. Blanket [Sha98c]. blankets [Tsi76]. blanks [NA92f]. Blanton [Han98a]. Blast [Ano61k, Hol04, Kap74, Ano94r, Rot93t]. blasts [Alb87b]. Blatant [Mie76]. Blechman [Mil86b, RK89a, Roc81]. bleeding [Obe97a]. Blessed [Gre87]. Blessings [Bor66]. Blew [Coo75b, Sch74b]. Blight [Leb94, Liv51]. Blind [Pra04a, Pra05a]. blinding [Ark95d, Her95a]. Blindsided [Pra04a]. blindspots [Fri88]. blinks [Isa92f]. bloated [Den14]. Bloc [Jon59, Hal93b]. block [Bar85b]. Blockchain [VU22, Dro22d, Pra22]. Blocking [Amb01, Ano86g]. blocks [Pow91b, Tsi77b]. Bloggers [Har06]. Blood [Uri94c, Tin23]. Bloody [Day77b, Nea04]. bloom [Ano93-27, Roe91, War49]. Bloomfield [You66c]. Blooming [Hod69]. blooms [Ano96m]. Blossoms [Edm74]. Bloustein [And65]. Blow [Hav03, Sim59c]. Blowing [Com76a, Rot04b, Ano93w, Tuc77, Ing79b]. blows [Ste97, Mes83a]. Blue [Ano93w, CJ01, Ano94v, Bin79, Thr92, Hul98]. Bluebeard [Mul73b]. Blueprint [RC70, Gar88d, Moo97o, Per92b, RD91]. Blueprints [AH92e]. Blues [CJ01, Ano91-41, Ark91, Cro18, Flo89a, Khr96b, NA88a, Isa96b]. bluff [Gar88c]. bluffing [Leb88b]. Blum [Moo95l]. Blumberg [Ber78a, Her90b]. Blume [Ano78l, Smi75c]. Blumenthal [Dug91, Van87]. Blundering [Got87]. Blurring [Pod16]. blurs [Hoe87]. Blyth [dK60]. Blythe [Lap57]. Board [Ano46-81, Ano17c, Bul08d, Bul08e, Bul08f, Ano96g, Ano16a, Cho95, Ano46-76, Ano63y, Ano77l, Ano89v, Ano95s, Gor51b, GM54, Han69, Law84b, Cou54]. Boardman [Ehr82a]. Boards [Ano48a]. Boas [Nor64]. Bob [Har88c, Sha91a, Dro24b, Isa96f, Sto17a]. Bobbled [Wur96b]. Bodansky [Pos79]. Bodega [Dav65, Hed65, HH65]. bodies [Ano95m]. Body [Jac74b, Koh76, Pre75, Ste78, Tuc08, Bar77b, Kel10, Lif76]. Boffey [Fay75]. Bogged [HI94]. Boggling [Hep77, Sla76]. Bohr [All56l, GH94, Haw94, Opp63d, Pet63, She86b, Swe02, Wil85b]. Bois [Sim87a]. Bold [Gne77]. Boley [Min94]. Bolshevik [Gan52, Shi51]. Bolt [Dav79b]. Bolté [Sto56b]. Bolton [Bos05]. bolts [O’S86]. Bomb [Alb89a, AH91a, AH92e, AH92d, AH93b, Alb94f, ABH02, AH03a, Alb06, ABB^{+}50, Als51, Ano45i, Ano46-98, Ano46f, Ano46i, Ano46h, Ano46k, Ano46g, Ano46j, Ano46x, Ano46s, Ano46t, Ano46-67, Ano46-71, Ano46-87, Ano46-84, Ano46-99, Ano46-107, Ano47l, Ano47-28, Ano49-40, Ano50m, Ano50-37, Ano50-46, Ano52i, Ano53u, Ano54i, Ano55r, Ano61o, Ano66f, Ano93m, Ano93-31, Arn50, Arn54b, Aro50, Bac50, Ber93b, BBSS50, Bet50, Bet63, Bet93, Bin10, Boy86b, Bro47a, Bro47b, Bro48, Bru47, Bur05, Can63, Coh00, Com46, Con60, Coo56, Cra67a, Dea54a, Dow83b, DH90, Dys61, Dys62, Dys63c, Ebe47, Far86, Fel71, For86, Fra83, Fri70a, Gav51, Gol88b, Gol95b, GR67, Gre87, Has57, Hay55, Heb51, Hei68, Hig48a, Hig66, HM90a, Hoo93, Hot76, Ing54a, Ing54b, Ing56c, Isa02a]. Bomb [Jan50, Jeh60a, Joh81a, Kay49, Kih47f, Kis54, Kna54, Kri02a, Kup06, Lal65a, Lan98, Lap48, Lap56b, Mac54a, Mar64b, Mar54c, MNS46, MHD^{+}05, Mer54b, Moo95d, Moo01e, Mor50b, Mor83b, Nar67, NA97c, Olk77, Ols62b, Opp51b, Opp68, Pas55b, POF49, Per00, Pot49, Que70a, Rab47b, Rab47i, Rab47l, Rab52d, Rab54d, Rab56a, Rab58d, Rab60g, Rab68g, Rho95, Rid50a, Rin61, Rot55, Rot89a, Ruf90, Rus46, San76a, Saw54a, Sch80c, Sha98a, Shi47b, SdSP62, Sho46, Sil71, Sim55, Sin05b, Smi58d, Smi63, SD56, SL46, Sti47, Sti48, Str54a, Str54b, STS^{+}49, Szi60a, Szi63b, Thi50, Thi55, Tru47, Vil75, Wal90a, Wea76, Wes58, Win98, Woo92, Wyl54, You66a, Zei53, Zem06, Alb87c, AH91b, AH91c, AH91h, AH91i, AH91g, AH91j, AH91k, AH92i, Alb95, Alv84]. bomb [Ano53o, Ano89m, Ano90-49, Ano91x, Ano97n, Ano15f, Ark97a, Bai95, Bar83b, Ber84c, BV98, BHM91, Boy85, Buk93, Cah76, Car03, CN91, CNB95, Coh87, De 05, DeW79, Dys20, Ebe20, Eli81, Fel77m, Fel81j, FR76, FS95, Fra90b, Gei16b, GP92a, GP92b, Gol84b, GBG^{+}81, Gue14, Gus04, Han95a, Hir89a, HM97, Hol87b, Hol94b, Kap81b, Kar14, Kha80, Kib17, Kir55, Kis78, Kno90, Lan89c, LS94, Lau91b, Len14, Mag85, Mag88a, Man84b, MSDP80, Mel97, Mic81, Mti96a, Nat84, Nei82b, NA90k, NA91b, NA92k, Nor02b, Nor03, NK86b, Oli89, OB78, Pab88, Pow93, Pre17, Rob78b, Rob78c, Rot85, Sar83, Sin17, SK84, Thi46, TFM81, Tsi83h, Ure50, Wei13, Wes71, Yem87, You76a, ZA04, Ber23, Dow77, Dow82h, Dow84c, Mia15a, Mos85, Wal90b, Wei48, Ano45c, Bet90]. bomb [LRD^{+}91, SB84, And83, Boy05, Cal62, Cof66, Day88, Dow82h, Gre64a, Gri83a, Han95a, Mar59a, May89, McC96, Mes87, Mil85a, Rab68g, Ram88b, Shi54c, Smi68, Som84b, Swe93, Wea82, Ano46d, Col54]. Bomb-Grade [Kup06, Sin05b]. bomb-making [Alb87c]. Bombardment [Mic55, Bro85b]. Bombed [NA88k]. Bomber [Blo05, Koc04, Pai82b, NA87i, NA90q]. Bombers [Ark99b]. Bombing [Ano46f, Ano49-40, Ano51q, KS58, She83a, Alp92, Loe85, MSW20, NA88g]. bombings [Cap98, Cog82]. Bombs [Alb87b, Ano47n, Ano53-41, Ano53-46, Ark00g, Bet54, Bon51, Bru57, Cap98, Cog82, Fer70, Har99, Ikl80, KC03, McC57, NA92m, NKH03a, Nor04, NK06e, Pug02, Rab55c, Rab56b, Rus50, Sar85, Sch75c, Smy64, Szi47a, Ano92w, Ano93r, Ark93a, Bea86b, Bel94, Ber87b, Che89, Coa47, Cum92a, Gou95, Hal93b, JL87, KK98b, KN14a, Lem81, Mak24, May90, Mos88, Nat77, NBM85, NA90k, NA97f, Nye13, PI92, Rot77, Sea85a, Sør77, Zac92, Zag95a, Say84, Bar90c]. Bombshell [Ing65, Hib95a]. bonanza [Isa93a]. Bonn [Mul69]. Bonner [Pla57b]. Bonnie [Sto18a]. Book [A.52, Abb84a, Abb85, Abb91, Abb93, Ack85, Ack87a, Ack95, Ack97, Ada52, Ada87, Ahe91, Ale69, Alf84, All55l, All55b, All55a, Alv87, Amr51, And83, And55, And78a, Ano45c, Ano48y, Ano49-30, Ano53e, Ano61b, Ano61c, Ano61d, Ano62k, Ano78e, Ano78d, Ano78f, Ano79a, Ano81c, Ano85d, Ano85b, Ano85c, Ano86k, Ano86h, Ano86l, Ano86j, Ano86i, Ano87p, Ano87r, Ano87o, Ano87n, Ano87m, Ano87l, Ano87q, Ano87s, Ano88d, Ano88e, Ano88c, Ano89f, Ano89g, Ano89e, Ano89d, Ano90p, Ano90n, Ano90m, Ano90o, Ano91q, Ano91s, Ano91r, Ano91-66, Ano93v, Ano93n, Ano94h, Ano98g, Ano98h, Ano99f, Ano01a, Ano05b, Ano10a, AM86b, Asp79, Aue78, Bal79c, Bal79b, Bal88, Bam85, Bar79b, Bar88, Bas80, Bei84, Bel93, Ber79, Ber84a, BD79]. Book [Bir87, Bla82a, Bla84, Bla85a, Bla87a, Ble87, Bou98b, Bou02, Bou87, Boy87, Boy89b, Boy91, Boy95, Boy05, Bra89, Bra87, Bre87, Bre93, Bri86, Bri89, Bri80b, Bro55a, Bro55b, Bro79b, Bro81, Bru50, Bru55a, Bur90c, Bus85, Cah94, Cal82, Cal84a, Cam54, Cam89, Cap98, Cha79a, Che88, Chi78, Cog82, Coh69, Col54, Col84b, Con81a, Con88, Coo81, Cop65, Cor94, Cra83a, Cra83b, Cra84b, Cra84d, Cra84e, Cra84c, Cra88, Cum98, Dal82, Dar91, Das96, Dav79b, Day75b, Day88, Day97, DeB81b, DeB81a, Der88, Deu83a, Deu83b, Dia83, Dia84, Div79, Diw79, Don86b, Don86a, Don89, Don90, Dow86, Dow88, Dow81a, Dre82b, Dug91, Eas85, Ede89, Ehr79, Ehr81, Ehr82b, Ehr82a, Ehr83, EE85, Ekl80, Eld92, Eli85, Els90, Eng91, Eng88, Eri95, Eva86]. Book [Fel81a, Fel83a, Fel84b, Fel86a, Fel92, Fel83d, Fel83e, Fin90, Fis55, Fis83, Fla90, Fly98a, For83, For87, For80, For86, Fra81, Fra82, Fra83, Fra91, Fre84b, Fre84a, Fre88, Fre95, G.84a, G.84b, Gai92, Gal86, Gan82, Gan83a, Gan83b, Gan83c, Gan84, Gar83a, Gar97a, Gen84a, Gen84b, Gen84c, Gen86, Gil83, Gil85, Gle94, Gol92a, Gol93a, Gol88a, Gol88b, Gol95a, Gol96a, Gol96b, Gol93d, Gol95d, Gol83c, Gol88g, Gom84a, Goo84, Goo86, Got87, Gov93, Gra93, Gra84b, Gre82b, Gre87, Gri83a, Gro78a, Gus83, Gus85, H.84a, H.84b, H.85, Haf79, Hal93a, Ham91c, Ham93, Ham83, Han90c, Han92b, Han94, Han95a, Han98a, Har85b, Har79b, Har81, Har86e, Har88c, Har92b, Har92a, Haw50, Her83a, Her85, Her88b, Her90a, Her90b, Her82b, Hir88, Hob85, Hoe88, Hol83a, Hol86b]. Book [Hol86a, Hol91, Hos87a, Hos87b, Hul98, Hun87, Hun85a, Hyd75, Hym91, Ing82, Ing79b, Ing79a, Iri79, Iri86, Ise92, Jac92, Jas81, Jer87, Joh97a, Jon87, Jon90b, Jor87, Kal68, Kap91a, Kar80, Kat52a, Kat52b, Kel98a, Kel79, Keu83, Key81, KFM50, Kil78, Kin80c, Kin80b, Kin86, Kli81, Kli90, Klo85, Kor92, Kra94, Kre83a, Kre85a, Kre90, Kre93, Kre89c, Kre98, Kri85, Kus83, LaF90a, LaF86, La 91, Lac85, LL90, Lan55, Lan56b, Lan98, Lap55a, Law88, LeR50a, Leb91a, Leb91b, Leb92, Leb94, Lei92, Lep89, Ler80a, Lev84b, Lin84a, Lin85, LP92, Lof51, Lon82, Lon85c, Low93, LD79, Lyn78, Lyt79, M.84, M.85, Mac84b, Mad79, Mag81, Man23, Mar79a, Mar79b, Mar69, Mar91a, Mar85a, Mar89e, Mar94, Mar97a, Mar98c, May89, McC88a, McC93, McC96, McE88, McG84, McM49a]. Book [Mea93, Mei52, Mei90, Men92b, Men91d, Mer79, Mer90, Mes82, Mes83a, Mes83b, Mes84b, Mes84a, Mes85a, Mes87, Met83, Mil85a, Mil88, Mil85c, Mil85b, Mil86b, Moo93a, Moo93b, Moo94a, Moo95d, Moo95b, Moo95c, Moo95e, Moo96b, Moo96c, Moo97a, Moo97b, Moo98b, Moo00a, Moo02d, Mor82a, Mor79, Mor85a, Mor83b, Mor84a, Mor87c, Mos92, Mos81b, Mos81a, Mos90, Mos91, Mye87a, Mye88, Mye89, Mye93b, Nad98a, Nad79, Nas82, Nat87, Nei82a, Nel84, Nor88, Nor02a, Nor04, Nor06, O’C87, Oze92, Pal02, Pap88, Par80, PMRP50, Par88b, Par86, Pat79a, Pat81a, Pat82a, Pat83, Pei80, Per91a, Per94, Per95, Per87b, Pit91, Pol93, Pol91b, Pos79, Pow88, Pow89b, Pow90a, Pow91b, Pre86, Pre96, Qua80, Rab48c, Rab98, Rad83, Ram88b, Ran85, Rau91, Ray81, Rea80, Rec89, Rep85a, Rep86, RK87a]. Book [RK89a, RK93, Rob91b, Roc81, Roc83a, Roc85, Rod85, Rom82, Rom89, Rom90, Ros90a, Rot79a, Rot89a, Rot93a, Rot96b, Rot97b, Rot97c, Rot98a, Row88, Sac84, Sag86, Sar85, Saw79, Saw80, Saw81b, Saw81a, Sca91, Sch90, Sch81a, Sch86, Sch69d, Sch79b, Sch80e, Scl79, Sco79, Sei81, Ses83, Sha90, Sha50, Sha84b, Sha91a, Sha91b, Sha81a, Sha95b, Sha98b, She79, She82, She83b, She83a, She89, Sho92b, Sie85, Sig85b, Sil79a, Sim84b, Sim87a, Sim85, Sim88, Sin81, Siv80, Sla88, Smi90a, Smi83, Smi85d, Smi88, Smi97, Sol80a, Som84b, Som91, Som84c, Soy89, Sta85, Bul05, SB91, Ste91a, Ste89e, Ste80a, Ste86c, Ste95, Sti83, Sto55a, Sto55c, Sto55b, Sto85, Stu89, Swe86, Swe88, Swe89a, Swe93, The89e, Tay87, Tay91, Tay98, Tel49, Tho80c, Tho80d, Tho80b, Tic94]. Book [Tir88, Tod78a, Tod79, Tod80, Tod81, Tra90, Tsi89a, Tuc97, Tum87, Uhl95a, Uhl95b, Uhl96, Uhl97a, Uhl97b, Uhl98a, Uhl98b, Val86, Van87, Van91, Van90b, Van03, Van88, Ven80, Wal81a, Wal85a, Wal87a, Wal83b, Wal84, War86, War92a, Wei85b, Wei89a, Wei89b, Wei48, Wei55b, Wel79, Whe84b, Whi89, Wig62b, Wil81a, Wil92a, Wil71, Wil80a, Wil80b, Wil84, Win84b, Win90, Win94, Win98, Win81a, Woo92, Woo82, Wou81, Wun82, Yar80a, You66a, You82, You84b, Yud83, Zag91a, Zel92, Zim81b, vL92, Ano82n, Ano91-68, Bro89a, Kee56, Ure56, Ano78n, Rot89a]. Booklet [Ano46d, Ano49h]. Books [A.54, Abr73, Abr74, Ada61, Ada67, Adl69, Alc63, All54a, All54b, All54d, All56l, All56k, All59l, And65, And68, And71, Ano49i, Ano50d, Ano50e, Ano50f, Ano51c, Ano52d, Ano53f, Ano53g, Ano53h, Ano53i, Ano54d, Ano54e, Ano55c, Ano56d, Ano56c, Ano56b, Ano57c, Ano57d, Ano57b, Ano60f, Ano61q, Ano62l, Ano66a, Ano68a, Ano72c, Ano78g, Arm68, Asi62a, Bar58a, Bar67a, Bar73a, Bar66a, Bar67b, Ber75a, Ber66a, Ber78a, Ber73c, Ber77b, Bet58, Bet65, Beu49, Bie68, BGR61, Bor52a, BB63, Bou61, Bre65, Bre73, Bre74, Bro57a, Bro60a, Bro57b, Bro47c, Bro54, Bro61, Bru57, Bru59, Bul67, Bur65, Cah76, Cal62, Can70, Cav57, Cav58, Cav61, Cha62, Cha58b, Cha74, Chr68, Cle74, Cof66, Con67, Cow65, Cre68, Cro65b, Cro67, Dam61, Dam62, Dan57, Dav69]. Books [Day75c, Day77b, DeN74a, Ded60, Ded62a, Dev69, Dew55, Din69, Div70, Dob54, Doh57, Doh58, Dra69, Eck78, Eis67, Eml66, Fal61b, Fay75, Fei77, Fel66b, Fel68b, Fel73a, Fel77o, Fer53, Fer54b, Fer57a, Fer57b, Fer62, Fer69a, Fin65, Fin57b, Fla68, Fol74, Fol75, For80, Fra61a, Fre76a, Fre76b, Fri64, Fri62, Fri70b, Fri72, Gab73, Gal54, Gar72, Gin54, Gol55, Gol78a, Gol67a, Gol69b, Gom71, Gor71, Gor73, Got64, Gou78b, Gra62, Gre54, Gre55a, Gre58, Gre64a, Gre67b, Gre68b, Gre71, Gro56b, Gru70b, Gun66, Hag77, Ham61a, Ham61b, Han75, Har60a, Har56, Har64a, Has71, Has57, Haw47, Haw56, Haw58b, Haw58a, Hec70, Hei54, Hen71a, Hig66, HI72, Hog68, Hol77, Hor77, How65a, How65b, Hug57, Hul76, Hul78, Hur78a, Hut61, Hyd75, Hym68, Hyn69, Ing56c, Ing70]. Books [Jac80b, Jen63, Joh51, Kal78, Kal63a, Kal62b, Kam71, Kap61, Kap67, dK60, Kat53a, Kat54d, Kat54a, Kat54f, Kat54c, Kat54e, Kat57a, Kat57b, Ker68, Ker54, Kin69, Kol63, Kon67, Kru74, LaF62, Lak66, Lal65c, Lal67b, Lal67a, Lal70, Lam69, Lan56a, Lan56b, Lan57, Lan58, Lan76, Lap54c, Lap56a, Lap57, Lap59a, Lap59b, Lap69, Las62, Lep62, Lew62a, Lew72a, Lew73a, Lis58, Lit54, Liv69, Lon78, Lyn69, Mac62, Mar72a, Mar78a, Mar58a, Mar59a, Mar54e, Mar54f, Mas66, Mat56, McK74, McL49, Mea62, Mei68, Mei77a, Mes64a, Mes64b, Mil69a, Mit54, Mol62a, Mor59a, Mor62b, Mor61a, Mor62d, Mot68, Mul65, Mul66, Mul72, Mul73b, Mun54, Nan73, Nas62, Nef78, Nel74a, Nic54a, Nie63, Nor64, O’B75, Ols59, Ols62a, Ols62b, Olu71, Opp62a, Ore58a, Ost72, Pal64b, Pan63, Pee78]. Books [Phe56, Phe63, Pla57a, Pla57b, Pla60, Pla64a, Pla69, Pog54, Pol68, Pom66, Pom67, Pon67a, PF75, Pre62, Pri70, Rab53a, Rab60b, Rab64a, Rab65c, Rab67a, Rab68a, Rab69b, Rab71a, Rab71b, Rab72a, Rab73e, Rab73d, Rab73c, Rap65a, Ras55, Ras63, Red69, Rei63, Rei78, Rei72, Rei74, Rei68b, Rit67, Rob56, Rob69, Rob62, Rod70, Ros68a, Ros67, Rot62, Rub73, Rus51a, Rus62b, Sac72b, Sak69, Sal63b, Sal71, Sar78, Sch54a, Sch69a, Sch71a, Sch78a, Sch68e, Sch68f, Sch70b, Sch63c, Sco71, Sco74, Sea72b, Sei69, Sei71b, Sei73, Sei60, Sei62, Sel63, Sha54b, Sha59, She60, She61, Shi50, Shu78, Shu49, Sid69, Sil71, Sim54a, Sim62, Sim65, Sin67a, Sin76, Sin78, Siv76, Smi54c, Smi54a, Smi54b, Smi58e, Smi58f, Smi59b, Smi59a, Smi60, Smi61]. Books [Smi68, Smi75c, Sne57, Sno68a, Sno68b, Sno69, Soc54, Sol68, Spe68, Spi66a, Spi72a, Spi75, Sta64, Ste76a, Ste77b, Sto56a, Sto56b, Sto58, Sto59a, Sto59b, Sto68a, Str67, Sut54, Tel76, Ten75, Tho62, Tod78b, Tor61, Tra71, Tsi72a, Tuc72, Tuc77, Ull63, Upt63, Val73, Vuc69, Wag61, War49, Was62, WW62, Why62, Wig62a, Wil53b, Wil68b, Wil69c, Wil69b, Wil74b, Wil76b, Wil77b, Wol62, Wol63, Wol55, Woo65, Woo73b, Wou73, Wri47a, Wri63a, Wri67, Wur68, Wyl56, Yon68, You56, You65a, You66a, You66c, You66b, You68a, You76a, Zac76, vH77a, Ano60h, Ano61f, Ano61g, Ano62n, Ano98l, Pri80, Tod78a, YES^{+}21]. Boom [Har91b, Opp04, Lan90m, Dow80b]. boomer [GM90]. Boomerang [HJ76]. Booming [Pra89a, Rot92f, Rot92u]. Booms [Ano70d, DeN65]. Boomtown [Lon98]. Boon [FP90a]. Boondoggle [Pra95]. Boorstin [Gen84c]. boost [GP18, Moo96d]. boost-phase [GP18]. Boosting [HGV10]. Boosts [AW85, Har90d]. Bootstrap [Hof61]. Border [Bet98, Woo00a, Woo00c]. Borders [Fly06b, Mül93, R.83, Rab73e]. Boris [Dar91, Ehr81, Sha94, Arb93]. Borklund [Kap67]. born [Azh95b, Ano56b, Bor65a, Bor65b, Bor65c, Bor68, Gre81, Hew81, Wil67a, Gre82b]. Borovik [Zag91a]. Bosch [dF79c]. Bosnia [Bos06a, Sha96, Sha97, Sho92c, Woo97a, Woo98a, Woo00b, Woo04]. Boston [Sch80e, Siv80]. Both [Gil83, Ano83b, BED93, Coo12, HSV19, MZ96, Plo86, Ung83]. Bottle [Kal61, Rab73b]. Bottom [Fit75, Yor83]. Bottoms [Isa93n]. Boulding [Was62]. Bouncing [PI92]. Bound [Men00]. boundaries [Cha79b]. Boundary [Cas77g, JOO15]. bouquet [Isa93m, Tho89]. Bouquets [Kas60].

Boutwell [Ano86i]. Bowden [Rei78]. bowed [Spa95c]. Bowles [Sto56b]. Box [Ise92, Ano63u]. Boxed [Woo03]. Boy [Nor04]. Boycott [Zhe90c, Ame80]. Boycotting [Tur78]. Boyd [Ano51-29]. Boyer [For86]. Boys [O’N94, RL89, Sut89, All59l, Boy95]. Bracken [AM86b, Mor84a]. Bradley [Beu49, Bra47, Bra49, Ham83]. Bragg [dF79h]. Brahmins [Wil69c]. Brain [Ded70, Fay75, Fre67, Hoa77, Joh65a, Jor67, Mor71, Mur70, Pra06a, Spe66, Sut68b, SG64, Whe84a, Cav90]. Brainard [Eav19]. brainer [Moo94h]. Brains [Mur68]. Branch [Ahe91, Ano87u, Ano92l]. Brandt [Bra87]. Brandwein [Har56]. Brass [Flo89b]. Brass-to-brass [Flo89b]. Brasstacks [Das96, Das96]. Brauch [Kre85a]. Braun [Ano85d, Neu07]. Brave [CLR06, Ano92w, Dro22c]. Bravo [Kas75, Ham91c]. Brawl [Woo00c, Woo00a]. Braxton [Wol58a]. Brazil [Alb89d, BMPd89, Buc77, Fly01a, Fly03, Gal76, Gol91c, Kra92, Pet96]. Brazilian [Lop78, Kru81, Mor14, San16]. breach [Hib95d, Van89]. Bread [Bor71, Han75, Har92b]. bread-and-butter [Har92b]. Break [Cic74, Fed50, FO07, Ler53, DPT17, Fre93a, Isa90g, Sco90, She87, Van91, Say84]. Break-Through [Cic74]. Breakdown [Pai83f]. Breaking [Abt84, BW84, Bla61b, Bro86a, Car83b, Day97, Fas78, FvHA86, Gay83, Gol69c, Hop84, Ing84, Kel84, Lev62, Lod84, Lon84, OG10, Pei84, Saa84, Sab63, Som83a, TT83, Tsi82a, Wie84, Yor83, ZG83, WC17]. breaks [Ber89a]. Breakthrough [Day75n, Ore60b, Rat71b, Rob75b, Lef77a, Cra88]. Breathe [CG04a]. breathing [Mor84c]. Breed [Fei77, Veg98b, Fei77]. Breeder [Abr74, Cha68b, CFM^{+}10, Fre76c, Haf50, Lew72d, Lov73, Men53, Pat86a, PR14, Smi74a, Ste77a, Fei77, Jan15, RR15, Tak17, Zha15a, vHJ97a, vHJ97b]. Breeders [Hec78, Pat80a]. Breeding [Dea53a, Zin53]. Brennan [Lap61a]. Brenner [Fra83]. Breslau [Ein48a]. Brettschneider [Krz20]. Brew [Ano68i]. Brewer [Kar80]. Brezhnev [Ano82k, Tum87, Gra82a, Kas75, Phy80]. Brian [Dow80e, Sal71, Sch80c, Krz20]. Brians [Hos87a]. Brice [Rez07a]. brick [Ano97i]. Bridge [Bas80, Fol60, Ols62a, Vie62, Odh79]. Bridges [Zam95d, Dow79d]. Bridging [Ard93b, DM18]. Bridgman [Hol62, Ure48b]. Brief [Ano50-30, Ano50-31, Ano50-32, Ano50-33, Ano50-34, Ano01i, Ano04g, Ano06h, Axe50, Axe51k, Axe51h, Axe51i, Axe51j, Bur09, Bur47, Ano50-35, Ano90-29, Ano90-31, Ano90-32, Ano90-34, Ano90-33, Ano90-30, Ano90-35, Ano91-33, Ano91-34, Ano91-35, Ano91-36, Ano91-37, Ano91-38, Ano91-39, Ano91-40, Ano91-41, Ano91-42, Ano92r, Ano92s, Ano92t, Ano92u, Ano92v, Ano92w, Ano92x, Ano92y, Ano93r, Ano93s, Ano93t, Ano93v, Ano93u, Ano93z, Ano93x, Ano93w, Ano93y, Ano93-27, Ano94j, Ano94k, Ano94m, Ano94n, Ano94l, Ano94o, Ano95i, Ano95j, Ano95k, Ano95n, Ano95l, Ano95m, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano96j, Ano97i, Ano97n, Ano97k, Ano97l, Ano97j, Ano97m, Ano98l, Ano98m, Ano98n, Ano98o, Ano98k, Ano98p, Ano99p, Ano99q, Ano99r, Ano00g, Bar76a, Bro79c, Hym91, Isa92i]. Briefe [Ano65e]. Briefers [Sho92c]. briefing [Ano91-68]. Briefs [Ano60h, Ano61b, Ano61c, Ano61f, Ano61g, Ano62n]. Brien [Rab52g].

Briggs [Phi84a, Mes83a, Par83]. Bright [Sim84d, Ano95k, OGFM18]. Brighter [Amr54, Bet58]. Brilliant [Her02, Moo92b, Opp90a, Ano90-34, Cha89a]. Brimming [Ano93y]. Bring [Ano90t, Ark97c, Ein50, RL89, Sut89, Yor75b]. Bringing [Ano94o, Mia98, Vof72]. Brink [Hal75b, Leb94]. brinkmanship [Elb81]. bristle [Bur87]. Britain [Ano57b, You76a, Ano91y, Ano01b, App46, Aur71e, Aur72b, Aur72c, Aur72e, Bri50, But04, Dew52, Dil57, Flo89a, Hin53, Joh04, Kin57, LB53, Men68, Mos93a, Mot48, Nel92c, NA90f, NA91b, NA93a, Opp02b, Opp03a, Opp03b, Pat90a, Pei70, Pie73, PI92, Pri53, Roy76, Rab47b, RD80, Sha92c, Ski53, Sut68a, Tuc77, You76a, Zim75, All54a, Mos81a]. British [And68, Bar73a, Han94, Smi60, A.55, Ali57, Als51, Ano46w, Ano46x, Ano46y, Ano46-30, Ano47d, Ano47e, Ano49k, Ano50-49, Ano51y, Ano53k, Ano54g, Ano57h, Ano58c, Ano60i, Ano68c, Ano99o, Ash50, Ash57, Att54, Aur72d, Cou48, Cal66, Chu54a, Coc51, Cun52, Dav52, Dew53, Dro20a, Erd52, Far71, Fil76, Fuc70, Gla56b, Gug52, Het55, Hod59, Ken87c, KN11a, Mas58, Mic51, NA87c, NA92b, NBF94, NA96a, NA99b, NA00b, NAKH01a, NK05a, NK13, Noy46a, Oli52, Orr56a, Pei50, Pei52, Pei85, Pic51, Pol52, Pri52, Rot88, Rot93i, Rus62a, San57, Shi47c, Shi47d, Sim53, Sim59a, Smi72a, Vic78, Wil87, Wri63a]. Britons [Atl47, Joh01]. Britt [McK23b]. Britten [Tra71]. Brittle [Nel84]. Broad [Bac58, Hol58, Kuz58, Osb58, Han92b, Pre86]. broadcast [Ack89a]. Broadcasting [Hav69, Pal77a]. Broader [Mul65, Smi75a]. broadsword [Rob78b]. broccoli [Ano98p]. Brockman [Ano78n]. Broda [Ano84f]. Brodeur [And78a, Jau90a, Jau90b, Sha90]. Brodie [She60, Por82]. Broglie [For80]. Broida [Fel81a]. Brokdorf [Gla12]. Broke [Pod05, Dun90]. Broken [Fri48a, Gro07, Plo89, San90, Zhe95, Sch75d]. Broker [SC07, Fel82c]. Bronk [Han76]. Bronowski [Bro69, Gro78a, Haw58a, Kat54a]. Brookhaven [Dow83c, Ano50-28, Sam86, Wol86]. brother [Rot92c, Ano98n, Joh94b]. Brown [Bal79c, Gre58, Han75, Hol87a, Leb91b, Nor06, Pla57b, Roc85, Tod79, Wou81, Gol93c, LG15, Ano85l, Ano14e, Ano14f, Bal79b, Ehr83, Gor71, Lan98, Mad79, Mes84b, Mit54, Mor62b, Rab73e, Sei69, Sim84b, Uhl97b, Wil68b]. Browning [FL96]. brownout [Arn96b]. Bruce [Abb84a, AM86b, Bri86, Bri89, Bus85, Dav79b, Leb94, Shu78, Smi83, Bul07b]. Brugioni [Leb92]. bruisin [NA88g]. Brunswick [Ano48d]. brush [DP12]. Brüyn [Bro55b]. Bryan [Wil53b]. Brzezinski [Her83a, Mul65]. Brzoska [Ano86i]. Bubbles [All59l, Day75i, Rot04b]. Buchan [Sei60]. buck [Fel76c, Gor96a, Mar58a, Smi59a]. Buckle [Bro20]. Bucks [Gla02, Col93b, NA88k, NA89j]. Bucky [Boe76]. Buddhist [Ahm00c]. Buddies [Moo01e, Ano91-34]. Budget [Ano52l, Ano64i, Ano89i, Dow83e, Isa90c, Isa94f, Isa02c, Isa05c, Joh67a, Mar65b, Mur84, Rot93d, Sto53, SG69c, Tam71b, Woo86, Ada83, BR90, CA87, Cai88, Cai89, Cai90, DW90b, Gre83a, Isa81b, Isa82d, Isa82a, Isa84a, Isa84b, Isa86e, Isa89l, Isa90z, Isa93i, Isa94g, Jon90a, Kor21, Lan90m, Mec19f, Mec21c, NA88n, NA88a, Ros89b, Rot92j, Ano86j]. Budgetary [Sco79]. Budgets [Ano68k, Eps75c, Jac74b, Wol63]. Buenos [Gol90d, Lem88]. buffer [Fis84]. Bug [Har06]. Bugs [Ano90u, KC03, Sie06a]. Build [FS08, Fis91, Mic68, Pai83b, Rab61l, Rab62i, Ure50, BW84, Bli76, Fie21b, Gol11, Haf84, Lem81, Lyn84, Mei20, Pai83f, Rev21, Ste85a, Tud21, WB85a]. build-down [BW84, Haf84, Lyn84, Pai83f, Ste85a, WB85a]. builder [NA87j, Mar89e]. Building [Ano95m, Dow77, Doy22, Gol95b, GW92, Hoo93, Nar67, Pow91b, Bul08c, Str06b, Swe79b, Tel76, Wat74, Bor82, Hol96a, HK21, Len11, Tsi77b, Wie84, dTdLF24, Dow83c, RK87a, Bar90c, Ehr83]. buildings [Lon52b]. Builds [May89]. Buildup [Pai82c, Ark85a, Ark89c, Mil81, NA89m, Pan81, Wri89]. Built [Ano46i]. Bulbs [AM74]. Bulgaria [Iva02, Pug78, Pav98a, Woo97b]. Bull [Dow81f, Har92a, Har92a]. Bullatomsci.org [Rot97d]. bullet [Hia87, Rot94g]. Bulletin [AZ89a, Alb93b, Alb93a, Alb94b, Ano68l, Ano75g, Ano82c, Ano82d, Ano83h, Ano83g, Ano83f, Ano89i, Ano89j, Ano89k, Ano89l, Ano89m, Ano90r, Ano90-29, Ano90-31, Ano90-32, Ano90s, Ano90t, Ano90-34, Ano90-33, Ano90u, Ano90-30, Ano90-35, Ano90w, Ano90x, Ano90y, Ano90v, Ano90z, Ano90-47, Ano90-27, Ano90-28, Ano90-36, Ano90-37, Ano90-38, Ano90-40, Ano90-39, Ano90-41, Ano90-42, Ano90-43, Ano90-44, Ano90-45, Ano90-46, Ano90-49, Ano90-48, Ano90-50, Ano90-51, Ano91v, Ano91-33, Ano91-34, Ano91x, Ano91-35, Ano91-36, Ano91-37, Ano91-38, Ano91-39, Ano91-40, Ano91-41, Ano91-42, Ano91y, Ano91w, Ano91-27, Ano91z, Ano91-28, Ano91-29, Ano91-30, Ano91-32, Ano91-46, Ano91-44, Ano91-45, Ano91-47, Ano91-50, Ano91-52, Ano91-51, Ano91-49, Ano91-54, Ano91-55, Ano91-53, Ano91-56, Ano91-31, Ano91-57, Ano91-58, Ano91-59, Ano91-48]. Bulletin [Ano92n, Ano92m, Ano92o, Ano92r, Ano92s, Ano92t, Ano92u, Ano92v, Ano92p, Ano92w, Ano92x, Ano92y, Ano92z, Ano92-27, Ano92-28, Ano92-29, Ano92q, Ano93r, Ano93p, Ano93s, Ano93t, Ano93v, Ano93u, Ano93z, Ano93x, Ano93w, Ano93y, Ano93-27, Ano93q, Ano93-28, Ano93-29, Ano94j, Ano94k, Ano94m, Ano94n, Ano94l, Ano94o, Ano95i, Ano95j, Ano95k, Ano95n, Ano95l, Ano95m, Ano96k, Ano96l, Ano96m, Ano96n, Ano96o, Ano96j, Ano96i, Ano96p, Ano96q, Ano97i, Ano97n, Ano97k, Ano97l, Ano97j, Ano97m, Ano97p, Ano97o, Ano98l, Ano98m, Ano98n, Ano98o, Ano98k, Ano98p, Ano14a, Ark89a, Ark91, Arn92a, Cav90, Col93b, Cor95, Day96, Eps90a, Eps93, Fly97, Fly98c, Fly98b, FR98, Gay98, Gol91b, Gor96a, Gor96b, Gor96c, Han90a, Han92a, Hib94b, Hib95a, Hib95b]. Bulletin [Hib95c, Hib95d, Hib96a, IM97, Isa83c, Isa89b, Isa89a, Isa89c, Isa89d, Isa89e, Isa89f, Isa89g, Isa90c, Isa90d, Isa90e, Isa90f, Isa90g, Isa90x, Isa90j, Isa90h, Isa90k, Isa90i, Isa90l, Isa90b, Isa90m, Isa90n, Isa90o, Isa90u, Isa90s, Isa90p, Isa90t, Isa90r, Isa90q, Isa90y, Isa90v, Isa90w, Isa90a, IG91, Isa91c, Isa91b, Isa91d, Isa91e, Isa91f, Isa91h, Isa91g, Isa91i, Isa91j, Isa91k, Isa91q, Isa91m, Isa91p, Isa91o, Isa91l, Isa91n, Isa92b, Isa92c, Isa92d, Isa92i, Isa92e, Isa92f, Isa92h, Isa92g, Isa92j, Isa92a, Isa92k, Isa93b, Isa93h, Isa93a, Isa93i, Isa93c, Isa93d, Isa93e, Isa93j, Isa93f, Isa93g, Isa93k, Isa93l, Isa94a, Isa94b, Isa94c, Joh93, Joh94b, Joh94c, Joh94d, Joh94e, Joh94f, Jon91a, Kal82a, Kal82b, Kri92b, Lan89a, Lan89b, Lan90c, Lan90e, Lan90d, Lan90f, Lan90g, Lan90h]. Bulletin [Lee91a, Loc89a, Low98, Mat97a, Mat97b, Mat98a, Mat98b, Mat97e, Moo94b, Moo96e, Moo97c, Moo97d, O’N95, Per90, Plo89, Pra89a, Pra89b, Qua80, Rot91a, Rot92d, Rot92j, Rot92e, Rot92f, Rot92g, Rot92h, Rot92i, Rot92k, Rot92l, Rot92m, Rot92n, Rot92o, Rot92p, Rot92q, Rot92u, Rot92t, Rot92s, Rot92r, Rot93c, Rot93b, Rot93i, Rot93d, Rot93g, Rot93f, Rot93h, Rot93j, Rot93k, Rot93l, Rot93m, Rot93n, Rot93p, Rot93o, Rot93q, Rot93r, Rot93t, Rot93s, Rot93e, Rot93u, Rot94b, Rot94c, Rot94d, Rot94e, Rot94g, Rot94f, Rot94h, Rot94i, Rot94j, Rot95d, Rot95b, Rot95c, Rot95e, Rot95f, Rot95g, Rot95h, Rot95i, Rot96c, Rot96d, Rot96l, Rot96e, Rot96f, Rot96g, Rot96h, Rot96j, Rot96i, Rot96k, Rot97e, Rot97g, Rot97h, Rot97f, Rot97i, Rot97j]. Bulletin [Rot98b, Rot98c, Rot98d, Rot98f, Rot98e, Rot98g, Sak82, Shk92, Sho89, Ste89a, TM97, Wal90e, Weg96a, Wur91a, Zam95a, Zhe90a, vH90a, AM86a, And10, Ano47k, Ano69d, Ano87u, Ano91-43, Ano91k, GR63, Mec15b, Nel75, Ano92-27, McC54c, AS62, Ack84, Ame68, Ano48r, Ano48s, Ano49j, Ano50g, Ano51k, Ano51d, Ano52e, Ano53j, Ano53q, Ano54c, Ano54f, Ano55d, Ano56e, Ano57e, Ano58b, Ano59b, Ano60g, Ano61e, Ano62m, Ano64h, Ano65d, Ano66a, Ano67c, Ano68b, Ano69b, Ano71a, Ano72d, Ano73a, Ano74a, Ano75b, Ano76a, Ano77k, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano78h, Ano78a, Ano78b, Ano79b, Ano80e, Ano81d, Ano82b, Ano83e, Ano83o, Ano84c, Ano85e, Ano86m, Ano86n, Ano87t, Ano87z]. Bulletin [Ano87-32, Ano88f, Ano89h, Ano89q, Ano89z, Ano89y, Ano90q, Ano90-56, Ano90-57, Ano90-62, Ano90e, Ano90f, Ano90h, Ano90c, Ano90g, Ano90b, Ano90d, Ano90i, Ano91t, Ano91u, Ano91e, Ano91f, Ano91d, Ano91b, Ano91g, Ano91h, Ano91c, Ano91a, Ano91i, Ano91l, Ano91j, Ano92k, Ano92l, Ano92i, Ano92g, Ano92j, Ano92c, Ano92d, Ano92a, Ano92b, Ano92e, Ano92f, Ano92h, Ano93o, Ano93a, Ano93e, Ano93f, Ano93d, Ano93h, Ano93b, Ano93i, Ano94i, Ano94g, Ano94c, Ano94d, Ano95h, Ano95f, Ano95b, Ano95g, Ano95d, Ano95c, Ano96h, Ano96-28, Ano96c, Ano96a, Ano96e, Ano96d, Ano96b, Ano97h, Ano97p, Ano97e, Ano97b, Ano97f, Ano97a, Ano97c, Ano97d, Ano98i, Ano98j, Ano98b, Ano98e, Ano98c, Ano98a, Ano98f, Ano98d, Art85]. Bulletin [BPR52, BFG^{+}85, BA89, Bro23, Com72, Dav55, Day74a, EH72a, Ewi60, Fer87, Har59, Jac72a, Jac73b, Jac73c, KW15, Ken62, KA88, Lee91a, Lew71a, Mec15a, Mec20e, RG49, Rab53b, Rab56d, Rob90a, Sal79b, Say92, Sco92, The89d, Tuf54, Vog87, Wes91]. Bulletins [Ano99g, Ano99h, Ano99i, Ano99j, Ano99k, Ano99l, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano04a, Ano04b, Ano04c, Ano04d, Ano04e, Ano04f, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano06a, Ano06b, Ano06c, Ano06d, Ano06e, Ano06f, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Kal82c]. Bullets [Siv75, BC16, Smi92b]. bullion [Isa90u]. bull’s [SC97]. bull’s-eye [SC97]. Bumbling [Lan89c]. bumpy [Bro20]. Bundestag [Mül89]. Bungles [Isa91j]. Bungling [Nas04b]. Bunker [Sch03, Sie04b]. Bunn [Kre93]. Bupp [Haf79]. Burbulis [Gut93b]. Burden [Ark00c, Mak01, Olt52, Ros52, Ano91-60, Dum86]. burdened [Fla87]. Burdens [Ste07c]. Burdick [Phe63]. Bureau [Ano45c, Ano53x, Sac46, Ano53t]. Bureaucracy [NH08a, Ing76, NH08b]. Bureaucratic [Cla74]. bureaucratization [Nas80]. Bureaucratized [Whe64]. bureaucrats [Rot94c]. Buren [Bas80]. Buried [Ing57a, Tuc01, Bun83]. buries [FLLW91a]. Burlazki [Leb91b]. Burma [Ahm03b, Kel10, Lar07]. burn [Dro22b]. Burnaby [Siv76]. Burnet [dF79g]. Burning [Koe22, KM18, Pie22, Rob56, AGB22, Ano93v, Dro22i, Gas93b, McK22]. Burns [Fin65, Som84c, Sho92c]. Burr [Han98a]. Burrowing [Ano95l, NA89f]. Burrows [Han94]. Burst [Hal75b]. Bursting [Day75i]. Burton [Ano12b]. Bury [LL97, Lym97, Eps77d, Hol77]. Bush [Ano85u, Ano98g, Cha84, Dia83, Ad 90, Ack89b, Ack90c, Ano52r, Ano92r, Flo89a, Isa89f, Isa89j, Isa91d, Isa92c, Isa92d, Isa92m, Isa93m, Isa01a, Isa04a, Isa04b, Lak66, Lea89, Moo91g, Pai04a, PMRP50, Pow04, Sha84d, Zac92, vL92]. Business [Ano49h, Ano92p, Aug50, Bel02, Bro57a, Lil47b, Mil85a, Wal69, Wes01, Ald76, Cha89b, Day75g, FP90a, Har89b, Hol94d, Loc95, Moo95c, Pra89a, Rot92e, Rot92f, Rot92u, Sch75e, Sut91, Wur93b, Ano78n, The89a, Ano48g]. businesses [Sto18b]. Bust [Opp04, Ano80o]. busy [Nor81, Zag91b]. Butler [Pla57a]. Butter [Ano93w, Str06e, Gol92a, Har92b, Vid90]. Button [Ark02d, Ano93x, Day76f, AM86b]. Buy [Lon75, Mac65b, Mil89b, Wei70a]. Buyer [Har90e, Hus90]. Buying [Pik96, Ros62, Wes04]. buys [Ano91-47, RD80]. Buzz [Rot92d]. BW [Ang60, Fot60, Rie60, Sim60c]. BWC [Rap11, KND10]. bygone [DW90b]. Bygones [Alv03, Alv20]. Byong [Uhl98b]. Byong-Moo [Uhl98b]. Byrd [Isa96a]. Byrne [Han98a]. bystander [Hil12]. bytes [Nye13].

C [All59l, Ano45j, Ano85d, Ano86k, Ano87q, Ano87-32, BGR61, Boy05, Bru50, Bul67, Bur65, Coh81, DM24a, Fer62, For80, Fra83, Haf79, Ham61a, Haw50, Kap67, Ker54, Kru74, Lam69, Lan55, Lap55a, LeR50a, Mar72a, Mar78a, Met83, Mil85a, Mye89, Nad98a, Ols59, Pla57b, Pla59a, Rab98, RK93, Roc81, She61, Sin67a, Sta64, Sti83, Swe93, Ure46c, Wri63a, Wyl56, dF79e, Fed68a, Spa95a]. C-17 [Spa95a]. C. [Ano48-38, Fal61a, Med47, MB48]. C.I.S. [Ano92-31, NA92c, NA92i, NA93d, NA93n, Nor94a, NA94d, NA95b, NA95d, NA96d]. Cabbage [Wea99]. cabbages [Woo97f]. Cabinet [Ano94n]. cables [Cla16]. Cadogan [Gro47a]. CAFE [Isa89m, Dow83e]. Cage [Ano90v]. Caging [Moo97b]. Cahill [Ano75e]. Cahn [She79]. Cairo [Jac64b]. Cake [Isa99a]. Cal [Ano94u]. Cal-1 [Ano94u]. Cal-3 [Ano94u]. Calamity [Aur72a]. Calculated [Joi54]. Calculating [LE85]. Calculations [Sta87, Lee24].

Calculator [Ano78n]. Calculus [Eva86]. Calder [Fer57a, Fel68b, Wag61, Ano57b]. Caldicott [Dow83b, Cal86a, Jr18, Dow82c]. Caldwell [Min94]. Calendar [Loc93c]. California [Ano46e, Ano14e, Che50, Cul14, DC14, Dvo65, Gar88a, Kre11, Lee74, Mec14, Mec23b, Tur14, War14, WM78, Wil82a, vdB65]. Call [Ano75c, Ano95j, Atl47, Bor73, Dow84b, Jac72b, Low98, Moo99a, Mur84, Rot03c, Wil53a, ASB^{+}84, Ano84k, Ano88u, Ano90-36, Ano91-36, Ano92y, Ano94n, Ano97m, Day77a, FE23, Gar78a, Hop84, HJS^{+}92, Mat83a, Nat86a, NA90q, Pra97, Wel80, Dow81d]. Callahan [Saw79]. Called [Ano46-104, Dil57, Szi60a, Woo01d, Ack94, KGS25, Wei80b]. Calling [Isa90d, Szi47b, Nei82a]. Calls [Ano46-27, Ano49-42, Bra47, Dro23c, Dro23h, Fra76a]. CalPERS [Dro21b]. calutrons [AH91f]. calve [Mer24]. Cambodia [Pra95]. Cambridge [Ano80b, Nor02a, Ano77l, BH53, Sal69]. Came [Lor01b, Moo95m, Pra05c, Rot00c, Woo55, Isa89l, Per20]. Camelot [Col06, Low66]. Camels [Wil75g]. Camera [Bri05, Leb92, Sch05b]. cameras [Dro22f]. Camo [Ano96k]. Camp [Ack97, Ano98n, Ano46-68]. Campaign [Ano61j, Dug91, Far08, Fel81e, Kal82c, KS58, Sin79b, Ber80, Cam79, Day96, ER84, Fre81, Gau24, Ial18, Isa78, Kre88c, Mec20h, Ano83g, CY81, Kal82b, Gom81a]. Campaigning [Got62]. Campbell [Gol92a, Wun82, Bro47c]. Camps [Ade02]. Campus [New62a]. campuses [Wal82]. Can [Abb12, Ack09, Aik61, And56, Ano58l, Ano90o, Bet46, Bor69, Bro72, Bru76, Bus47, CAM20, Coo12, Dal92a, Dan72, Deu60, Dje72a, Dro22b, Eav17a, Ein50, Eis66, Fel79a, Fre72b, Gib91a, GLSS07, Gou75b, Gus69, Hen93, Hib01, Hir72, Hoa72, Jac74d, JH16, Kah59, Kat90a, Kin67, Kin84b, Kre88b, Lil46, Loo50, Lym08, Mac65b, MR21, Mec20b, Mec21f, Mec22c, Mel58, Mic68, Mor87c, Mot49, Mot51, Nye13, Rab62h, Rab72b, Rad92a, Ram72, Rev66, Rot61, Rus58b, Saw63, Say84, Sch02a, Sho80, Spo51, SK03, Ste69f, Szi50a, Thi55, Tsi75a, Uhl97a, WSG^{+}15, Web06, Wei66b, Wes79a, Ano91-69, Ano97n, Cla77, Coo78, Dou24, Dro23e, Dro23i, Fai97, Fal79, FTvHW76a, Fie19, Fie21a, Fra79, Gro98, HK21, Int81, Isa92n, IS15, Kal81, Kap80, Kas77, Kib12, Kis79b]. can [Kor17, Kre88c, Lam83, LS92, Lan79, Lin86, Lor92, Lym96, Mac92, McK23a, Mec21b, Mec23b, Meh21, Moo25, Nay12, Pie22, Pow89a, Rus20, Sal79a, SMĆ08, SD09, Smi85a, Squ12, Tro20, Tud21, Vad21, Zag91b, Zha12, Zha19, ZG83, dTdLF24, vD78, Bet20, Ken60b, Mil85b, Yar80a]. Can-Do [Web06]. Canada [Ano96k, Ark86c, Ark89d, BM22, Bra02, CO03, Eps87a, Hoc57, Jac80a, Lau47, Mar61b, McN48, Pug81, R.83, RL96, RL89, Sha10, Sim87b, Smi69a, Spu72, Tru84, Zei85, Ber79, Dew55]. Canadian [Ano87n, RI83, Tho84a, Ano46z, Ano91-42, Car84a, Dou15, Eps87a, Hue12, Lam87, Lew50, Mar90, McC69, New67, Sha87a, Sim87b, Sok87, Sok91, Tru83]. Canan [Deu83b]. canary [Cha82a, vH77a]. cancellation [Wit52a, Wit52b]. Cancer [Ano78k, Eps77b, Lac57, Mor78, Oft71, Red69, Sac72a, Wil92a, Ano77j, CMR79, Hil12, MK80, Ste90b, dF79c, Ano78k]. cancer-risk [Hil12]. candidates [Bel20b, Hec16, Mec16e, Tho80f]. Candor [Ano55e, Lap54b, Lap55b, Lil53, Rab53f]. CANDU [Maw78a]. Canning [MB99]. cannon [Die91]. cannons [Bla20]. Cannot [Adl57, CS16, Moo25, Sap80b]. Can’t [Cum02, Rab51h, RK93, Ano83b, Ano92r, Dro23c, Isa89m, MZ96, Sch24, Ung83, Woo98c]. Cantelon [Sti83]. Canyon [Mos67]. Cap [LH70, Cle11, Zam92c, Ano98l]. cap-and-trade [Cle11]. Capabilities [Zah69, CAH84, KK98a, KN18b, Len21, Cal84a, Mar94, Ano84g]. Capability [Fri74a, Kul88, NA87i]. capable [NA87b, NA87l, NA88p]. Capacities [Maw78a]. Capacity [Com74, Mac54b, ML75, Maw78b, Net75, Sea65, Alb87c, Cot22]. capercailly [Ano96j]. Capers [Ano95l, Ano93w, Ano94j, Ano94m, Ano94n, Ano96k, Ano96j, Ano97m]. Capita [Gol90b]. capital [DP20]. capitalism [Ano92u, Gol91b]. capitalist [Dro24d]. Capitalists [Wei88]. Capitol [Ano74d, Mar64f]. capsule [IM97]. Captive [Fer54b]. Car [Cic74, Jon71, Mil71, Net69, Ano97i]. Carbon [Ano73g, Doc07, Dro22i, Gre08, Boo22, Cav13, Cle11, Cul14, Fie19, FS81, Gun20, JP13, Kam20, Mar82b, RF21, Sto19a, WMRK79]. carbon-14 [Mar82b]. carbon-free [RF21]. Card [Sch07b]. Cardinal [Woo92]. Cards [Fel68c]. care [Hia87, Kap76]. Careers [Mac66, Orl84]. careful [Bun82]. carefully [Rot91b]. Careless [Ing70]. cares [Tra90]. Carl [All60g, Bar79b, Ber75a, Con81a, Mar89e, Rob91b, Smi85d, Smi97, Sta79]. Carlo [Sch79b]. Carlos [Dro25a, Gor96c]. Carlton [Nei82a, Sch79b]. Carlucci [Cai88]. Carnegie [Whe84b, FCC^{+}46]. Carnesale [Wal85a]. Caroe [dF79h]. Carole [Pol93]. Caron [Ark88a]. Carpenter [Jac80b]. Carpet [Cre63]. carriers [Ano93k, Gif13]. carries [Joh94a]. Carroll [Sch78a]. Carrots [Gro07]. Carry [MV04, Ano95m, Fei90, Gei16b]. Carrying [Kis81, G.65]. Cars [Fer69b, HG71, Rus70a, SEM70, Wou71, Tur14]. Carson [Ehr79]. Cart [Sap75b]. Cartan [For80]. Cartel [Mar78c, Man77]. Carter [BWS^{+}80, For87, Ant78, BL79, Buc79, Cav79, Fel81f, Gol79, Gue79, Hol77, Kin78a, Law88, Phy80, Sig98, Swe80, Wil78a]. Cartwright [HD82]. Case [Ano53x, Ano86i, Art68, BST02, Bea89, Ber49, Edi54, Day75b, Dow83b, Dys69a, Gar91, Gou47a, Gro56b, Hur60, Kal54, Kal63b, Ke72, Keu83, Kra07, Lip54, Lov73, Lov74c, Mar64g, Mar58b, Mar53c, Moo99b, Mor87c, Mur70, Nat67, Nel74b, Osg60, Par47, Pol49, Pri69, Rad53, Rob75b, Rus58a, Rus62a, Sal64a, Sch69a, Scu65, Sho76, Sil64b, Sim54b, Wes58, Zin53, Ato54a, AP88, Alb94b, Ano90-40, Ano94k, Bad78, Cas77a, Col79b, Cro86, Dah77, Den82, Dic77, Eps77c, Gre77, Gre23, Han79, Hol20, JFS20, Joh83c, Lar87, Leb88a, Leg92b, Lei82a, Nat85a, O’B80, Pon17, Por19, RF21, SL13, San79, Sha86, Sha78a, Sin79a, Sok87, SG69a, SG71, Tar23, Tur98, Wag88, WG82, Wri97, Phe56, TR80]. Cases [Ano48t, Cle48b, Dav48, Sci55b, Sag86]. Cash [Ano95m, Ano90-30]. Cash-and-carry [Ano95m]. Cask [JvH09, RL14]. Casper [Dow82f, Mor82a]. Cassidy [Swe93, Why62]. Cassini [Rot96c]. Cassiopeia [Wil68b]. cast [Tur17]. Casting [Moo92c]. Castle [Mul73b]. castoff [FP90b]. Castro [Leb94, Men93b, Ark87g, Pea92a]. casual [Ano86o]. Casualties [Dea04, Abr83]. Casualty [Ano51-27, Bru47, Can63, Ano77n, Isa94e, Sic96c]. Catalog [For80]. catalogue [Ano93z]. Catalyst [Nel57, Ste76b]. Catastrophe [Dud75b, Hud60, Lar65, Per87b, Ind14, AH22, Bet76, DL78]. Catastrophic [Man23]. Catch [Ano90w, Tho77, Ano88n, Ano94j]. Catch-22 [Ano90w, Ano88n, Ano94j]. catching [Fin84]. catchphrase [Mec21b]. Catechism [SG75]. Catherine [Fin90, Sco74]. Cathleen [RK89a]. Catholic [Pfe52, Pol91b, Mil83a, Mor57a, Pop52, Pep15, Sie52, Tay52]. Catholicism [Mor53a]. Catholics [Fox88]. Cats [Rao77]. cattle [DeW83]. Catton [Ehr81]. Caught [Gan03, Obe98]. Caulfield [Fin90]. causation [Gre12]. Cause [Ano50x, Day74d, Day75i, Fly00e, Kih47f, Sim62]. Caused [Cog82]. Causes [Alb75, All54b, Har81, Ano94n, HS96]. Causus [Sch06a]. Caution [Dys59, Hol94c, Bar81d, Isa94d, Ids81]. Caveat [Cav05, Gol73a]. Cavendish [Gol82b]. Cayley [WW62]. CB [Ano66r]. CBS [Dow81d]. CBW [Bar80e, Bel72, Col00, Kar72, Veg98a]. CD [Mar62, Mil62, Wil62]. Cease [Bet95]. Cecil [Rab70a]. Celebration [Opp79a]. celebrities [Kra15]. cell [Dro22f]. Cells [Dob54]. Cellular [Mat18]. Censored [Ano88g]. Censoring [Ben01c]. Censorship [Ano91-34, Cle48c, Dud77a, New62a, Wer73, Zag95a, vH93, KK85, Par85, Yev86, Klo85]. Census [Ano52-27]. cent [Ano95m, Pen98]. Centenary [For80, HC12]. Center [Ano46-68, Ano96-27, For80, Kre90, Law11, Rig65, Saa84, Sal65, Spe72, Wil73a, Ano84n, Rab85]. Centers [Dje68, Dow92]. Centowski [Dow82g]. Central [Asp79, Don66, LP92, Mos84, Pra02, Spi60, Wio84, Ack88i, Lei78, Coa84, Gom83, Ken83, La 83, SL83, Woo00c]. Centre [Bar82a]. Centrifuge [AH04a, AH91i, AH92b]. Centuries [LH70, Ste65a]. Century [Aro56, Fra83, Gla76c, Got87, Lap73a, Lor00b, Man72, Pal02, Shr70, Smi90a, Wad74, CK20, Dan16, Deu84a, DM24c, Dic01, Dic03, Dic11, Eps92a, Kas77, Mec22c, Moo97j, Sim95, Tal18, Ezr72, Har10, Joh65c, Mor62d, Rip68, Bul04a, TS01, Ano98g, Bam85, How65b, Mar98c, Moo02d, Ste91a]. ceramic [GMV^{+}24]. CERN [Aug56a, Aur72f, Bak60, Day77e, Kow55, LT70, Ano68e]. Certain [Rap65a]. Certainty [Pos71, Wal77, Pos24, Ste83a]. Cesium [Tam74]. Cesium-136 [Tam74]. Cessation [Ano46-84, Fed59a, Fed59b, SD56]. Cetron [PF75]. CFCs [RC90]. CFE [FLLW91c, Isa91p, Sha95c, Web93]. Chacun [Gra81a]. Chadwick [Gol82b, Oli82, Lan98]. Chafer [Cal82]. Chagula [Ano78l]. Chaim [Ano85d, Cor53]. Chain [And73, Dea53a, H.85, Pla61, Sko58, Wat74, Bro79a, DuB62b, Eis62, Fre77b, Ruf90, Smi13]. Chained [Bel55]. Chairman [Ano52y, RS52, RDS^{+}52, BBG^{+}85, All58a, Ano46-27, Ano52b, Kel79, Phy80, Str54b]. Chairs [Rot92k, Joh95]. Chalk [Cal82]. Challenge [Ano60f, Ano86k, AM74, Aur72a, Bac03, Bor70, Bro54, Cam89, Car89, Dew48, Dom98, Fre60, Hun87, Ing56c, Kam71, Kan69, Kol63, Mar58d, Mar63, Mil88, Mit54, Moo97b, New58, Opp47a, Rab56g, Rab66e, Rub08, Sch76c, Sea70a, Smi71, Tel58, Whe64, Bre87, DMJ98, Fel77l, Fis86a, Gaf86, Hum77, Kas77, KN14d, LBCA23, Not15, Par17, Sig89, Sim95, Bur90c, Rab69b, Rub73]. Challenged [Tay69]. Challenger} [LN86, Wal86]. Challenges [NL61, Rab68f, Rab73f, War63, Ano98p, Bug22, DM24a, Isa86c, KSS17]. Challenging [Kol75]. chameleon [Ara94b]. champion [Arn97a]. Chance [Ing62, Moo02b, Rab56j, BRB16, Gro81, Moo92k, Tho83]. Chances [Coh60, Gra60, Arn47]. Chandler [Roc85, Dow84c]. Chandrasekhar [Ano95w, Cha70a]. Chang [Kre98]. Change [Ano80k, Ano81c, Ark01d, Bal79c, Ben58, Bet75a, Cal65, Cir07b, Dev69, Dow80e, Gru70b, Han66, Jac74e, Kal70, MF63, Mes84b, Moo94c, Myr74, Olu66, Pol67, Rob69, Rot01h, Sch08a, Bul08a, Ste61, Tae51, War68, Yar80a, Ald17, Ano90-34, Ano93z, Ano10c, Ano14b, Ano14e, Ano15c, Ano15d, Ano15e, AH22, Bac12, Bra13a, Bru15, Brz12, BW18, BKV13, BR15, CR17, CM12, DC14, Dha16, DS15, Dro21b, Dro22b, Dro23f, Dro24e, Eps12, Eze16, Gif13, GG16, Gle24, Gra16, HK18, HNR12, Hue12, Huq11, Huq16, Kas89, KGS25, Kre11, Krz20, MR21, Man23, MD18, Mec17b, Mec19a, Mec19d, Mec20c, Mec20d, Mez12, Mil12b, Moo18, Mor12, Myn11, Ols16, Pie22, RH17, Sal16, Sch18, Sch21, Smi13, Smi21, Soc15, Soc20, Som20, Squ17a, SW16, Uri94b]. change [Van20, Wan16, ZS15, Ehr81, Pla57b, Ros68a]. change-where [Dro23f]. Changed [Bro56c, Eis66, Rab56i, SS73, Gar89]. Changes [Ano49a, Joh65a, Mul55a, Rum71a, Nas03a, Ark95d, Bye88, Dro20a, Dro24g, Hol16]. Changing [Ast67, Aur71b, Bro57c, Bro56c, Fel83d, Flo89b, Lal66i, Mul47, Nor74, Orl83, Sco74, Sto59a, Jor18, Kun23, LG15, Mcc87, Pug77, RL78b, YG77a, Mor62d]. Channels [Ano87v]. chant [Loc92b]. Chaos [Ade02, Hau68, Dum93, Har90a, Sap88]. Chap [Lut73]. Chapman [Ano78n, Nor64]. Chapter [Fry57a, Pow81]. Character [Fow58, Lal65c, dJ57b, Dro23k]. Charade [Fel75b]. Charge [Dow81i, Ack87d, NA92j]. Charged [Gar78b, HWG95]. Charged-particle [Gar78b]. Charges [Ano48t, Ano53-40, Gol84a, Nic54b, Nic23]. Charging [Dro23a, Mec22b]. Chari [Das96, Per95]. Chariot [Boy95, O’N89a, Wid61h]. charisma [Pea92a]. charity [Han92a]. Charles [Bou61, Bro79b, Che88, Fel84b, For87, Gre54, Gro56b, Pla69, Rab71b, Rob56, Sch81a, Sto56b, Sto58, dF79h]. Charlie [Moo94i, Wei94b]. Charlotte [Lan56a]. Chart [Ano49-39]. Charter [Fah91, Wri47b]. chartered [Pea91]. chase [Isa93j, Sie05g, Fri72]. chases [Flo89a]. Chasing [Com75h, Moo91a, NA88b, Spa95b]. chatbot [Gou23]. cheap [Ano90-31, Isa92k, Mon89]. Cheaper [Ano91-40, MB99]. Cheaters [DMMH02, PD89, Wei91b]. Cheating [Kep82, Ack86, Mar83e]. Chechnya [Zag95c]. Check [Khr05, Moo98c, Pra04d, NA93i, Sco94a, Sep23]. checkered [Ram16]. Checking [Ack89c, Gay98]. Checkout [Ste69b]. Checks [Isa04c]. Cheema [Das96]. cheerleading [Hol94e]. cheers [TM97]. Cheery [Moo96m]. cheese [BV21]. Chemical [Alb54, Ano45m, Ano46-27, Ano46-88, Ano48d, Ano49n, Ano65g, AK01, Axt73, Ben01b, Ble54, Bur90c, Cas02, Cla61, Coo71, Dub60, Fed64, Fei90, Gla60, Gol85a, Hai54, Har71, Isa91r, Isa96d, JC54, Kap90b, KT97, Kel98a, Kyl97, Lap73a, Lun79, Mer54a, Mes64a, Mes70, Mie74, vCM89a, vCM89b, Rab54c, Rab59i, Rab60h, Rob75b, Rum54, Sei71b, Smi91, Smi92a, Smi92d, Smi93d, Spe49, Tho80g, Tod55, Tuc97, Tuc01, Tur73, Whi51a, WS95, Wri89, Ano68l, BBP14, Bur90a, Bur90b, DeV92, E.77, Fre91c, Gao17, Hog97, Isa88f, Isa89f, Isa90a, Mei98, Mes91, Rot92l, WF86, Wes85, Gar68, Har14b, Lan14, Sai14, Sid69, Ano46-56, Ano54y, Bur90c, Sei69]. Chemical/biological [Wri89]. Chemically [Ure55]. Chemicals [Smi93b, Zan91, Fre91a]. chemischen [Ano68l]. Chemist [Ano54p, Arn73]. Chemistry [Ber70a, Bur47, Hen68, Lan56b, Lan58, Ram70, Tuc97, Ano96l, Isa90b, Mec20f, Tuc10b, Ano87-32, McK51]. Chemists [Rot95d]. Cheonan [Sig10]. Chernobyl [Abr86a, Abr86b, Alb86, Ame86, Ano86y, Ano91-34, Ano96r, Bab93, Ber86a, Bet86, BM93, DH86, Fis86b, Fly06a, Fri11, Gal90a, Gal90b, Gal91, Gor11, Gor20, HS97, Hes97, Hof86, Isa90e, Leg87, Mar89d, MR90, Mar93a, Mar93b, Mar98a, McC86b, McC86a, McC88a, MM11, Mye92, Pat86c, Pat86b, Pet11, Sch11, Sch16, Set16, Sim16, Sta97, Tho86a, Tur87, Wei86b, Wei86a, WvH86, vHC86, Mar91a, McC88a]. Chernobyls [Bru96]. chested [Mag88b]. Chester [Gom71, Sto56b]. Chet [Ano78n, Lap58]. Chewing [NA87d]. Chiapas [Fly00b, Nad98b, Nad98c]. Chiarelli [Mil85c]. chic [Ano96k]. Chicago [Cam54, Wil71, AAF^{+}52, Ano45b, Ano45c, Ano46-65, Bla69a, Bur50, CD48, DuB69a, Mil69c, Mil70, Opp63e, Rab46a, Smi58d, Vil75, Wal69]. chicken [Ark85e, Mol96, Sei84]. Chief [Art71, Obu89, Rab48a, Sha54b, Mec24, Ano75a, Ano85l]. chiefs [Ano93z]. Child [Dje74, MM01, Wes97]. Child-Bearing [Dje74]. Childhood [Moo92d]. Childish [Hen17]. Children [McN63b, Mos92, Say84, Ano92r, Ano94t, Bri88, Dys81, McC96, Rob62]. Chile [Dvo65, Gol90e]. Chilean [CR17]. Chilling [Moo98d]. Chimera [KJE16]. China [Dow79d, Kus83, LL16, Sch70b, Adi69, AH20, Ano65e, Ano66b, Ano66k, Ano69c, Ano69d, Ano69i, Ano69j, Ano69q, Ano69r, Ano79c, Ano80j, Ark01e, Bet69a, BBZ01, Bin04, Bin18, Brz12, Buc50, Cha69a, Col93b, CS17, Cot22, Cum99a, Dal22, DeN76a, Dem19, Der69, Dro22a, Dro23g, Dro24a, Esp72a, Esp72b, Fel71, Fer61, Fie91, FS84, Fin84, Fra66, Fri69, Gao17, GvH23, Gom61b, Gun66, Hal66a, Hal66b, Han17, HB69, Hsu79, Hym71, Isa91b, Ke72, Kul17, Lac69, Lal64c, Lal66i, Lam80, LL12, LL16, Mar64b, Mar71a, Mec22c, Mec22b, Mit50, Moo99a, Moo99b, Mor65, Mur72, Nea63, NA90f, Oli66, Orl83, Orl84, Pai85f, Pan22, Pra89c, Raj17, Rot72, Sah20, Seg13, Sha66, She00, SIF66, Sim84c, SL80, Smi77a, Smi82a, Sok22, Sub68, Sut70, Sut84, TS22, TWSM61]. China [TW66, Vad21, Wan96, Wan83, WKK23, Wil67c, Wil69a, Wu22, Wu23, Yah69, Yan82, You64b, Zha99a, Zha12, Zha15b, Zha20, Zha19, Zhe95, Zho79, Pow88, Rab65c, Rab71a, Cof66, Iri79, Kli90, May89, Mye93b, Tra71, Uhl98b]. Chinese [Han94, Sak69, Sim88, Ano67d, Ano99m, Bin15, BR97, Cof65a, DE23, Dje74, Fel66c, Gal91, Gar77a, Hou72, Ing65, Joh97b, Kal70, KK19b, KN11b, KN13a, KN15a, KN16a, KN18a, KK20a, KK21a, KKR23a, KKJK23a, KKJK25, Lal65a, LA98, McA65, Myr66, NA87c, NA89d, NA89g, NA92k, NA93b, NBF94, NA96a, NA99a, NA00a, NAKH01b, NK03a, NC06, NK08a, NK10a, Pol79b, Pol09a, Ric76, Riq15, She94, Wan16, Wil78c, Yin15, Zha15a, Zha15c, Sim85]. chintzy [Lea89]. chips [GB77, NA87h]. Chloe [Wil68b]. chocolate [Zag94b]. Choice [Ano63r, Kim87, Lei68, Mot47, Rab60b, Tel73, Wei66c, Abb98, BPTB76a, BPTB76b, Fel80a, Fos95, Mye90, Nit80, NB77, Pav98b, Var96a]. Choices [Bas80, Gur88a, Her83a, Hul76, Mes84a, Rep86, Wes99, Woo92, Cai88, Hou82, KK18a, Bas80, Hol77]. Chomsky [Cro17d]. Choose [Spi75]. Chooses [Aur72b]. Choosing [Kar88b, NA87g, Pat81a]. Chop [Bet89a]. Chorover [Tho80g]. Chorus [Sch01b]. Christian [Ano94h, Tho62, Boe52, Com51, Dav57a, Koc57, McC56]. Christians [ZS15]. Christie [Mec19a]. Christmas [Ano97q, Coh87, Rof97, Sul58b]. Christopher [Ano78n, Bam85, Dow80g, Fis83, Lac85, Mes64b, Rea80, Roc85, Wal81a, Wou81, Mec23a]. Christy [Pon67a]. chromosomes [BBL13]. Chronic [DM18]. Chronicle [You56, Dar91]. Chronicles [Bro60a, Fin90, Rot98a]. Chronological [Ano50h]. Chronology [Ano51t, Saw54a, Sch73, Ame86, Ano92-33, Ark88a]. Chruschtchow [Leb91b]. Chu [Dro16a, Dro23a]. Chuck [Mar97a]. Church [Fly98a, Sch57a, Pop52, Pep15, Sie52, Bul10, Tay52, Pfe52]. Churches [Har79a, See79]. Churchill [Boy87, Ber87b, Ber87c, Chu50, Har86a, JL87]. Churn [Goo00]. CIA [Bou98b, Har88c, LP92, Ano94m, Har88c, Hol80b, Hor66, Kre87b, Mec16b, Pra93b]. cigar [Lut89]. Cigarette [Cha77]. cinematic [Mec19h]. Cioc [RK89a]. Circle [Bha99, Led71b, Ste91a, Dow84b, Com72, EH72a]. Circles [Bol63, Tay96, Pai81b]. Circular [Ano95k, Mol89]. circulation [Kri14]. circumvents [Con86]. Circus [Dav57b, Isa88d]. Cis [BB22, NA92d]. Cis-lunar [BB22]. Citation [Eis54a]. Cities [Ano46f, Aug48b, Aug48a, Lin55, Liv51, MTK46, ME51, Szi61b, Wol01a, Woo02, vH00, BBC^{+}98, Hau17, Jac15, Wol01b, Ano95l, Ash81, LD79, Mea62, Fri64]. Citizen [Abb91, All62b, Ano72g, Ano87s, Gen71, Had56, McD62, Rot92a, Wea57, Yar80a, Zac50, Kri78, Mec23b, Sin79a, Ano85d]. Citizens [Abr73, Ano66c, Fos88, Ray80, Sin79a, Wit03a, Ano77l, HGF17, KW15, Sak98, Sob21, Tho80e]. Citizenship [WW50]. City [Ano53n, Ano84k, Bet69b, Ede04, Lap48, Mar65i, Mon53b, Nan73, Ano77l, Eav19, KDFS81, Ato50, Ede20, Mar64c]. city-level [Eav19]. Civil [Ass63, Ano49l, Ano50i, Ano50j, Ano50k, Ano50h, Ano50-44, Ano51e, Ano51f, Ano51g, Ano51k, Ano51n, Ano62o, Ano62p, Ano87w, Ant50, Asp79, Bon49, Bon50, Cha53, Dee59, Exe57, Fed62, Fel57, Foz61d, Gai69, GL48, Hal76b, Hin69b, Hoh75, Hop49, Hut50, Hut51, Jaf51a, Jaf51b, Kna63, Kup51, Lap50, Lap54d, Lap58, Lap59c, Lap60b, Lar50, Lev52, Lev53, Lew61a, Mar66a, MFC62, Mau50, Myr69, New62b, Nor64, Pie62, Pit64, Rab50a, Roz64, Sim56, Sle52, Sle53a, Sle53b, Sle53c, Sle53d, Sle53e, Sle53f, Sle53g, Sle54a, Sle54b, Sle54c, Sle54d, Smi53, Sto53, Sym50, Wig66, Win84a, Wit03a, Zei50a, Zei50b, Zei52, Zei54, Coc81a, Fel77k, Kap78b, Kap78c, Lea87, Pat80d, Sie80, Ada52, Bri50, Ske90, Zei51, Mac62]. Civilian [Abb85, Abr83, Ano46-45, Ano46-60, Ano46-63, Ano48-37, Ano50-41, Ano54x, Ano78n, Hig48a, Mic55, Rab48e, Tay87, Zuc52, Abb90, Blu23, Bro87c, Hoe87, Kis89, Lan90e, Lov13, Rep85b, Squ13, Tsi90]. Civilian-Based [Tay87]. Civilization [Bor73, Bou62, Cha75c, Mor73, Pug88, vHW77, Pow91b]. civilizational [All15b]. civilizations [McK25a]. Civilizing [Fer69a]. Claasen [Leb91b]. Claiborne [Hol94d]. Claim [Ano07i, Ano07q, Cir05b, Gar97c]. Claims [Kwo10, Lei92, Mar93b]. Clams [Swe78a]. clandestine [Kla88]. Clarence [Ano78n, Sea72b]. Clarfield [Abb85]. Clarification [Ano48r, Ano52c, Kau80, Lym97, KH85, vH76b]. clarified [Lev90]. Clark [Ano78n, Hym91, Lak66, Mil85c, Sha59, Cal62, Cla66, Pee78, Shi50]. Clarke [Sid69, Kin80b, Kre85a, Lam69]. Clash [Isa89l, Tay62]. Class [Lon85c, Wei86c]. Classes [Ame69]. Classical [Smi61]. Classification [Fel55]. Classified [Fed67b, Sch68b, Ano91-37, Ano91-45, Gre81, Hew81]. classroom [BRB16, Fei81a]. Claude [Haf79]. clause [Loe87]. Clausen [Ham93]. Claussen [Sto18b]. Clay [Lap54c]. Clayton [Smi83, Tho80d]. Clean [Alb93b, Ano73h, Ano90x, Ano93r, Hal91, Pra05c, Woo00f, AK95a, Dro21d, Lop91, Mag88a, Rot95k]. Cleaner [Wei70a]. Cleaning [Ano98o, Rau92, Wol65, Ano88i]. Cleanly [Fot91b]. Cleansing [Rab49b, Nad98b]. Cleanup [Alv05, Moo05b, RMS03, VS04a, Ack97, Ano91y, Nel93b, Nil14, Rot95k, SM90, SM14, Ste91b, Ano91-63]. Clear [Ack89d, Ano96k, McN76, Rab47e, Fel80a, Gle91, NA93j]. Clear-eyed [Ack89d, Ano96k]. Clearance [Ano49b, Ano55a, Ano55i, Con55, Gre64b, Pub55, Sci48, Spi49, Swa49, Eva54]. Cleared [Ano48w, Ano49-28]. Clearing [Col94, Fly05b, Wur93a]. clearly [CS17]. Clemens [Kru74, You66c]. Clementine [Ano98l]. Clennam [Com48b, The48]. Clerical [Wei75]. Clever [Oli58]. Cliché [Doo59, Wea85]. Clichés [KD59]. clicks [Arn22]. cliffhanger [Sco84b]. Clifford [Tho80g, Haw58b, Hym91]. Climate [Ano81c, Ano14e, AH22, Bor69, Bra13a, BRB16, Brz12, CKA10, CM12, Dha16, DS15, Dro23f, Edw12, Fle70, Huq16, KGS25, Lan61a, LH70, Mec19d, Mec20c, Mec21g, MRH16, Sch08a, Sch21, SF81a, Smi13, Soc15, Bul08a, Bul09e, SW16, Sue61, vV23, AGB22, Ald17, Ano10c, Ano12f, Ano13e, Ano14b, Ano15c, Ano15d, Ano15e, Bac12, Bis14, Bru15, BW18, BKV13, BR15, CYP14, CR17, CAM20, CO14, Cor78, DC14, DR14, DP20, Dro21b, Dro22b, Dro23m, Dro24e, Dro24g, Eav19, Ema21, Eze16, Fie18, Fig23, Gif13, GG16, Gle24, Gra16, Hen17, HK18, Hue12, Huq11, Jor18, Kre11, Krz20, MR21, Man21, Man23, MD18, McK23a, McK23b, Mec17b, Mec19a, Mec20d, Mer24, Mez12, Mil12b, Moo18, Mor14, Mor12, Myn11, NLS25, Nor11, Ols16, PFW18]. climate [Pie19, Pie22, Rev21, RT12, RH17, Sah20, Sal16, Sca11, Sch19, Sch15b, Sez20, Smi21, Soc20, Som20, Spr19, Spr25, Squ17a, Squ21a, SMRVS22, Sto17a, Sto17b, Sto18b, Sto19b, Tro20, Van20, Wan16, War14, ZS15, Dro23m]. climate-change [Ano10c, Sal16]. Climates [dF79e]. Climatic [Rum71a, Ehr84]. Climatologist [Ano15a]. Climbing [Isa03a]. Clinging [Isa90-27]. Clings [Isa93f]. Clinton [Ano46i, Ano46-28, Ano92r, Ano95m, Dor96, Fos95, Isa92c, Isa93i, Isa94a, Kib04, Loc93c, NH08b]. Clintonizing [Sha93a]. Cloak [Nas06]. Clock [Ack89e, Ano69e, Ano72i, Ano06g, Ano07g, Cam60, Day74e, Fel82d, Hub60, Lau60, Phi84b, Rya89, Bul07c, Whi75, Ano90i, Ano91j, Ano98m, Ark96d, Dav55, Fel78i, HD92, HJS^{+}92, Lyt85, Moo96a, Sal79b, Spa96, Sue82, Ano91k, Dro23i, Kra10, PF82, SS24, Bul02d]. Clockwatching [Nel75]. Cloning. [Tho80g]. Close [Cho03, Eas00, Eme60, Lut89, NA90q, AH91c, Ark86c, Bra13b, Dro23c, Dro23h, FE23, Moo97l, Sez20]. Closed [Dua78, Mil69a, Str06a, Han95c, Leb88a]. Closer [BF05a, DeV88, Fel80b, Fel81d, Har10]. closest [DP12]. Closet [Har72b, Isa04e, AH92i]. Closing [Ack01, Ano92r, Com72, CJLT24, EH72a, HI05a, HI05b, IG91, Sot67, Tak17, Ise92]. Closure [Ack01, Isa93h]. cloth [Ack97]. clothing [Dri93]. cloths [Que81]. Cloud [Abb93, Ano15a, Tho80c, LG15, Mil86a, NA92e, Tsi83i, Ano83m]. Clouding [Cha75b]. clouds [Pra98, Dow80i, McE88]. cloudy [BRB16]. Club [Aik61, Cal60b, DeV62, Han75, Hol01, Kha98, Arn96a, Far77, Gan83e, Tho81a, Yan82, Cha73b, Kap74]. clubbier [Arn96a]. clueless [Eav16]. Clutter [Dav05]. Co [Chi69, Con46b, Tuc97, Ark92, Mec22c, Bar81d, Ids81]. co-conspirators [Ark92]. co-exist [Mec22c]. Co-operation [Chi69, Con46b]. Co. [Chu93]. Coal [Fri54, Jac73c, Sin78, Ano11h, Pie22, Bas80, Dow80d]. Coal-Driven [Fri54]. Coalition [Ano90y, Pom77]. Coalitions [Dea91]. coast [Sin79a]. Coastal [Lee74, Jac15]. COAT [vHW76b]. Cobalt [Arn50, Gei16b]. Coca [Hod00]. Cocaine [LP92]. Cochran [Cal84a, Abr74]. Cockburn [Lei92, Lei92]. Cockcroft [Ber66a]. Cocked [NA87e]. Cockroaches [Rot89a]. Cocoyoc [Ano75c]. Code [Bon90, Day97, Mos80, Seg13]. Coercion [CL00, Cot22]. Coexistence [Fed70, Kno55b, Kol63, Nea60, Pon60, Rab68c, Hui91]. Coffey [Gar72]. Coffin [Cof62a, Cof62b, Cof62c, Cof62d]. Cognitive [DH98]. Cohen [Ano90o, Das96, Par86, Ack85, Bus85, Moo94b, Mor83b]. cohesion [Sha82]. Cohesive [Bar66a]. Coin [Smi74d]. Coit [Win84b]. Coke [Ano91-35]. Col [Lap55a]. Cold [Ano90-30, Bou98b, Bou02, Boy87, Boy91, Bus85, Dow81f, Dug91, Fel77o, Fel80c, Fly99c, Gar97a, Gol93a, Gra62, Gra93, Gus83, Ise92, Joh64c, Kra94, LaF62, Lov89, Moo97a, Mul65, Nol17, Pri86b, Sha91b, Sim87a, Smi85d, Uhl95a, Van03, Ano91-36, E^{+}84, Isa90j, Kin78a, MK02a, Mul80, Nin88, Sch92a, Tay89, Wes79a, Ano90i, Ano91j, Ano91k, Ano92t, Ano97m, Ano13a, Arb77, Bra95, Cra86b, Cum92b, Cum97b, Dal22, Fel93, Fis62a, FLLW91a, Har86a, Hol96a, Isa90-27, Isa94e, Jac82b, Joh99a, Kal88, Kal95a, Mar89a, Mar02c, McD85, Moo92e, Nat55, Nel57, NK16, Off02, Pat87, Pei53, Pra92, Ros81, Rub08, Sch19, Sei50, SW21, TS22, Tsi89c, Uhl97b, Whe90, Wie67, Wu22, Bus85]. Cole [Tuc97, Kel98a, McE88]. Colglazier [Met83]. Colin [Hob85, Roc83a, Wou81]. Collaboration [Fed70, EF78, Joh51, SB89, TPPB21]. collaborator [Bor89]. Collaborators [The89b]. Collapse [Bet81b, Han90d, Spr19, Leb94]. collapsing [Mec20b]. Collateral [Blu23, Pra89b]. Colleagues [Ano57h, Git74]. Collected [Cha74, Gab73, Phe56]. Collections [For80]. Collective [Ber49, Den77, Pol91a, FR97a]. College [Gen86, Ano46-86]. Collide [Hol07]. collider [Cro86, Pri86a]. Collier [S.80]. Collina [Mec21d]. Collision [Joh81b, Kol75, Far98]. Colonel [Rec89]. Colonialism [Gou75b, Pou78]. colonies [Ano98o, Tho80d]. Colorado [Ano46-90, Ano81b]. Colton [Cra84e]. Columbia [Lyn69]. Columbian [Hym69a]. Columbus [O’L77]. Combat [Ste95, Nol17, Pai82i, YES^{+}21, Lap55a]. Combating [Uhl96]. Combats [Hil65a]. combination [Sha18]. Come [Fel75b, Flo90, Pra98, Sti68, STS^{+}49, AK95a, Bru76, Deu84b, Isa98b, Kas77, Kla82, Lac90, Mec14, Rot93u, Rot98c, Deu83b, Ano86h]. Comes [Alb93b, Fry57b, Lal66i, Mol69, Sie05c, Boo22, Cha88b, Dro22h, Wil77b, Fel68b]. comet [Ano85y]. cometh [NA89c]. Comfort [Ano51h, Cho03, Eas00, Ret99a, Ark86c, Kor92]. Comforts [Har79b]. comic [Ano98l]. Comics [Ano97m]. Coming [Ano93v, Ano95o, Gho11, Gor96d, Hen71a, Hor14, Les63, Lib59a, Moo03b, Nad98a, Rab71a, RMS03, Smi83, Ald76, Ben76a, Ben76b, Cor96, Isa91m, Mat18]. Command [Moo01d, Mor84a, NA87f, AM86b, Wur91a, Bai92, Smi59a]. Commanders [Sha91a]. Commandments [Bet72, Sch72d]. commandos [NA93m]. Comment [Ano46-89, Ano46-88, Ano51x, Ano53m, Ano53l, Ano54i, Ano57g, Ano69f, Ber72b, Bor58c, BKSG60, Col54, Com54, CCMU46, CFU47, Coo74, DP53, Dee57, Eis59, F.61, Fel61, Gil69a, Gol70a, GLMS60, Har72a, HG71, Hux59a, Kan71a, Kih47f, LB53, Lap59g, Lew71b, Lil53, Mar59b, McC56, McM49a, MCC51, Mil50, Mur54b, Nor47, Oga68, POF49, Pin57, Rab47c, Rab48d, Rab70b, Rap65a, RKR55, Rot58b, Rus55a, Rus70a, S.46, S.55b, SSW^{+}55, Sch71b, Sha54a, Sim54b, Sim55, Sim59b, Sim59c, Sim59e, Sim59d, Sim60b, Sim60a, Smi58c, Smi58a, Smi58b, Ste60b, Szi47c, Wer56, WB51, Ano86q, Gus82, Fel81b, Fel81c, Gib82, Gom81a]. Commentary [Ass80, AK81, Agn82a, Alb78, Alc81, AR80, Alv80, Ame80, Ano78i, Ano79d, Ano80g, Ano80f, Arn80, AL79b, Asp83, Atk78, Aus81, Avi83, Bal78, BØ79, Bar83b, Bar80c, Bar80d, Bar80e, Bar80j, Bar80f, Bar80g, Bar80h, Bar80l, Bar80i, Bar81b, Bar81c, Bar81d, Bar81e, Bar81f, Bar81h, Bar81g, Bar82a, BW80, Bec79, BW83, BT82, Bel02, BK81, Ben83, Ber82, Ber83, BL83a, BWH80, Bey80a, Blu84a, Blu84b, Bon83, Bor82, BF05a, Bre84, BL81, Bro79a, BR79, Bro80b, Bro84b, Buc79, BWS^{+}80, Buc82, Buk03, BCB81, But84, CW81, CT78, Cam79, Car81, Cha84, CFG^{+}81, Cha83b, Cha75b, Chi82, Cho03, CY81, Coc81a, CMR79, CB79, Col84a, Col79b, Coo79, Cop80, Cor84, Cot84, Cou81, Cra82a, Dav79a, Dav83, DD79, Dav79c, Dav82, DLF79, DeW81].

Commentary [DM81b, DM84, DG79, Deu84a, Deu84b, Dom82, Don82, DWH83, Dow78, Dre84b, Dud75a, Dun83, Dun49, EV84, Ege83, EF78, Eis84, Eli81, Eng78, EBC83, Eps79, Eps81, Eva84, Eva82, Fai79, FBH81, FH83, Fas78, Fei81b, Fei81c, Fel80a, FM82, Fis81, Flo81, Fre81, Fri82, FS81, Fri84, Fro49, Gal81a, Gal81b, Gei81, Ger84, GS81, GF79, Goe80, Goe79, Gol84b, Gol79, Gom80, Gom81b, Gom84b, GJS80, Gor80, Got83, Gra81a, Gra81b, Gra84a, Gre84b, Gre05, Gre83b, Gri83b, GC83, Gro82, Gro84a, GBG^{+}81, HP80, Haf84, Hai81, Hal84, Han79, Har79a, Har84c, Har79c, Har83, Hay81, Haz82, Hen78a, Hen84, Her79, Her78b, Her83b, Hir81, Hit03, Hod81b, HF80, Hol05b, Hol05a, Hoo81, Hud78, HC83, Hum83, Hun84, HNB81, HQL84, Int81, Ids81, IPB84, IES79, Isa81b]. Commentary [Isa81a, Jac82a, Jac80a, Jac82d, Jac83, Jan81a, Jan80, Jan81b, JM83, Joh80a, Joh81b, Jos81, Kal81, Kap81b, Kap83c, Kap79, Kat80a, Kau80, Kau82, Ken80, Kep82, Key80, Kin78a, Kin80a, Kin81a, Kin81b, KR84, Kin84b, Kis79a, Kis81, Kle81, Kol81, Kol84, Kom80, LaB81, LaM79, Lag80, Lag81, Lak78, Lak82, Lan79, Lan83b, Lar84, Lat79, Lat84, Law84a, Law84b, Leb79, Lei83a, Ler80b, Ler81, Lev80, Lew80a, Lew80b, Lin82, Lin79, Lof84, LLS83, Lug05, Lyn84, Mem79, Mem80, Mac83, Mac84a, Mac81b, MP78, Mac81c, Mal81, Man84b, MM82, MSDP80, MK80, Mas83, Mas84, Mat83b, Mat80a, Mat82b, Mat84, Mat80b, Mat79, Maz81, McA84, McC81, McC84, McK83, McM82, MW83, Med81, MW81, Mel78, MK79, Mic81, Mie79, Mil83b, Mil84, Mil78, MA83]. Commentary [MJ81, Mit83, Mon84, MD81, Moo03a, Mor80, MJ80, Mos80, Mot83, Mul80, Mus80, Nag81, Nai81, Nea80, Ne’81, NW79, NF83, Nit80, Nor81, Nor82, Nus80, O’B79, O’B80, O’B82, OGP84, Off84, Ola82, Old81, Osb82, Osg80, OS81, Pug78, Pug81, Pai81a, Pai82b, Pai02, Pan84, Par83, Pat78a, Pat78b, Pat79b, Pat79c, Pat79d, Pat79e, Pat79f, Pat80a, Pat80b, Pat80c, Pat80d, Pat80e, Pat81b, Pat81c, Pay84, Pen84, PF82, PP83, PG79, Pet81, Phi84b, Pie80, Pla78, Pla84, Pou78, Pre80, Pri81a, Pri80, PG84, Que81, Qui84, R.83, Rab05b, Ran82, Ras80, RGA81, Red78a, Red82, RI83, Ric79, Ric84, Roc83b, RS83, Röl80, Ros81, RR81b, Rot84, RAB^{+}80, Sak79, Sal80a, Sap80b, Sap80a, Sap84, Sax49, SG83, Sch84a, Sch80a, Sch81b, Sch81c]. Commentary [Sch80b, Sch79a, Sei84, Sha84a, Sha80, She80, She05a, SR81a, Sho80, Sil79b, SWC^{+}84, SD83, Sim84e, Sin83, Sin79b, SL80, Sle84, Smi80, Smi81, SL83, Smi84c, Sno81, SP82, Sol80c, Sol79, SR81b, SM81, Som84a, SF81b, Som83b, Som84d, Sør80, SS80, Sox80, SA82, Spe81, Spu80, Sta80, STRS80, Ste80b, Ste81b, Ste82b, SK84, Sto79, SSR83, Str84, Sue82, Sut83, Swe78a, Swe79a, Swe80, Swe81, Tay81, Tho80a, Tho81a, Tho84a, Tho80f, Tho80h, Tho81b, Tho84b, Tho84c, Thu79, TFM81, Tos82, TC83, Tsi83a, Tsi83b, Tsi84a, Van80, Ven79, WG82, WN79, Wal81b, WM82, Wei78a, Wei78c, Whe84a, Whi79, Wie83, Wil82a, Wil78c, Wil80c, Wil82c, WK82, WP80, Win79, Win81b, Wio79a, Wio79b, Wol78, Wol82, WM84, Woo81, WMRK79, Wri49, Wri84, Yan82]. Commentary [Yar80b, Yar80c, You78a, You81, You84a, Zha05, Zil83b, Zim79, Zwe84, dC78, dKL83, dV78, vHWF^{+}81, KFM50, Wei86d]. Comments [ACC^{+}67, Ano52u, Ano60e, Ano66f, Ano67e, Ano67f, Ano68d, Ano68e, Ano68f, Ano69g, Ash48, Bar61b, Bar67c, Bla67, BN67, Bor68, Bre68, Bre69a, Cha68a, Cof62a, Cof62b, Cof62c, Cof62d, Cof67, CCMU46, Cra67b, Cro68, Dow67, DN67, DB67, Dys61, Dys69a, Dys69d, Fed67a, Fed61, Fel62a, Fel67c, Fel67a, Fel68c, FDC52, Fre67, GLH^{+}79, Ger67, GJ67, Gor67a, Gre67e, Gre67d, Har70, Hau68, Hib68, HY67, Jen70, Kat61a, Kea61, Ken61c, Kih47f, Law69, Lei68, Ler48, Lew68c, Lew69b, Lew70f, MY67, Mar61c, MS61, Mas67, MH67, McG68, McM48a, Mun67a, Mun67b, Myr72, Nar67, Nat67, New67, Ney69, Ogb48, Opp51a, Pet61, Que67, Rab62d, Rab66b, Rab66a, Rab67d, Rab67c, Rab67e, Rab70b, Rea68, Ros68b, Rot61, Rub62, Sag69, Sch68a, SCT^{+}79, Sen67, Ste69c]. Comments [Ste69d, Sti67, Sto61, Stu69b, Sue68, Tam69, Tel48a, Thi58, Top58a, Ure48b, Ure67, UM67, Wei61, Wei69b, WF68, Wil68a, Wil67a, Wil68c, Wol68, Wor69, Wu60, Yat69, Yor70, You68b, vB67, Wil82b, Art71]. Commerce [MF10]. Commercial [BL83c, KM89, Kre89b, Lin88, Ric89, Spe89a, Nel12, Rot77, Ste89c]. Commie [Ano98n]. Commissar [Con88]. Commission [Ano46-62, Ano52b, Ano55y, Ano80k, Gre62, Jas81, Lan80, Lie46, Nie63, Soc55, Zer13, Den14, Fou84, Wur91b, Ano46j, Ano46-100, Ano46-104, Ano46-101, Ano46-102, Ano46-103, Ano46p, Ano46-43, Ano46-66, Ano46-74, Ano47c, Ano47j, Ano47r, Ano47z, Ano48i, Ano48w, Ano48-40, Ano48-28, Ano48-29, Ano48-39, Ano48-41, Ano49-28, Ano49-39, Ano50t, Ano50-35, Ano50-39, Ano51l, Ano54u, Ano55r, Ano82p, BAB^{+}46, Bey46, Bla46, BBD^{+}46, Bru47, Buc82, Can63, Gre80, Haf51, Ing52c, Kal54, Kal63b, Rab47d, Rab49a, Rab49c, RS52, Smi93c, Som84b, Sto19c, Swe78a, Tel48b, Van46, War48, Scl79]. Commissioner [Ano58h, Ano49m, Pik48]. Commissioners [Ano47u]. Commit [Cra01, KK18a]. Commitment [Bri49, Liv63, Sal80c, Sca11]. Commitments [Moo97g]. Committee [Ano45c, Ano45h, Ano46-38, Ano46-63, Ano46-94, Ano46-99, Ano47h, Ano47t, Ano47u, Ano50u, Ano51l, Ano53w, Ano55m, Ano55v, Ano56k, Ano56n, Ano57a, Ano70a, Ano85t, Bea47a, Bet75b, BBG^{+}85, Cog82, Kre93, Lev51b, Lon50, Mec21c, Pri53, RG49, RDS^{+}52, Spi50, Spi49, Upt63, vN57, Ano96s, Mül89, Ano46-29, Ano46-31, Ano46-57, Ano46-61, Ano46-65, Ano46-82, Ano48u, Ano48-36, Ano49l, Ano49z, Ano49-36, Ano51r, Ano52a, Ano53a, Ano53b, Ano53c, Ano53d, Ano53-42, Ano63k, Ano66l, Ano74i, Ano85w, Ano96-27, Bac58, BCH^{+}48, Ein49, Err58, Fer68, HPL^{+}49, Jac74e, OBC^{+}49, Rus51a, Sci48, Shi54d, Sim56, Tho80b]. Committees [Ano50-48, Isa86a, Isa85c]. Common [Ano63v, Ano78n, Aro55, Bel46, Cha70b, Gol55, Has68, Haw58b, Hum55, Kat54a, Mak90, Mat83b, Mor62d, Nac65, Pon67a, Rab60a, Sea63a, Swe89a, Tod64a, Whi76b, Yon68, ZGD83, Cle11, Eis84, Fin77b, Her60, Hol86d, Joh91, Kis93b, Lei83b, Pug86, Rot92b, Mac97a]. Commoner [But73, MW72]. commonplace [Bar80g]. Commons [Chu54b, Chu54c, Shi46a]. Commonsense [Mey66]. Communicable [Kap60]. communicate [Ano94l]. Communicating [Ano11c, Dav53]. Communication [Fri56, How56, Mey56, MD66, Rip68, Swa66, Ano91-39, Cha83a, Ano83b]. Communications [Alb71, Ame69, ADH^{+}73, Ano48v, Ano70b, Ano73b, Ano74b, Bro70, Bar73b, Bel72, Bet72, Bha70, Bon72, Bra74, Bro69, Bru72, Bur72, But73, CHP71, Cae73, Can72, CSG71, Cha72, CH71, Cow71, Cra72, Cre74, Dav72, Dje73, Dru72, Ele57, EH72b, Ell72, Elm71, Elm74, EB73b, Fes73, For70, Gad72, Ger72, Gol72b, GB72, Gom70, Gou74, Gre72, Haa73, HW73, Har71, Hen71b, Hol73, HG71, Hud70, Hym69b, Jai73, Jan73, Jon71, Kan71b, Kat73, Kor72, Kra72, KP71, Kug69, Lan70, Led72, Lei73, Lei70, Lev72b, Lev73, LI70, Lew71f, Lon74, Lut73, Mag73, Man72, Mar72b, McA71, Mer72, Mey69, MW72, Mor70a, Mou73, Mul71, Ner69, NW70, Obe74, Oft71, PTNR70, Pas74, PS72, Pay70, Pie61, Pie73, Pol73, Por71, Por73, Rab67b, Ric71]. Communications [Rig70, Ros69, Rot72, Rum71a, Rum71b, Rus71, Ryb71, SEM70, Sag71, SS69, Sch67a, Sch72a, Sch72b, She74b, Shu72b, Shu72a, Sib73, Sku72, Sme74, Smi53, Smi72b, Smy61, Spe74, Spi72c, Spi72b, SB73, Ste70, Ste72a, Ste73a, Ste72b, Str72, Sut70, Szi71, Tam73b, Tay74, Tri72, Wat70, Wat73b, Wee73, WC74, WZSA73, WI71, Wer73, Wes71, Whi72, Wol71a, Woo73a, You71, You73, Yut69a, Zei70, ZSA73a, ZSA73b, ZSA73c, ZSA74, vA73, vH72, Dor84, Mow95a, NA87f]. Communiques [Dou69, Lew68b, Lew68a, Smi69a]. Communism [Ano94k, Dou48, Ken53, Mar00, Mor53a, Sch70b, Smi60]. Communist [Bal88, Ano54p]. Communists [Zag95e, Zag95c]. Communities [Bam85, Moo51]. Community [Ano55c, Ano67m, Ano69l, Ano69k, Ano86p, Ćos65, Dub59, Fel69b, Gou78b, Han69, IES79, Kis69, Kri86, Mar48a, McE52, Opp59b, Pla53, Rab69a, Rab58a, Rab63g, Shi54a, UM67, Wyl56, Zei50a, Bro79e, Eng78, Hes78, Hun04, Ito14, Lyn86, Rot18, Mos91, Wil81a]. companies [Ano86q, BW18, Eav17a]. Company [Ano51l, Ben69, Gil69b, Gil69a, Net71]. comparative [Ric76]. Compared [Ano66n, Maw78b, Ano88s, Fel14]. Comparison [Ano47e, Ano60j, Ano62q, HE76, Ano86e]. Compartmentalization [Gol95b]. Compelling [Ber63]. compensated [Ano91-46]. Compensating [NA88c]. Compensation [Ano48b]. Competition [Cra84b, Kim22, Pla59b, Por71, Sin78, Ber87a, Cla16, Spa95c, Squ21b, Ano88e]. Competitive [Bec55, Mat97e, Sch77b]. Compilation [Cog82]. Compiled [Bre93, For80]. Complaints [Klo85, HD92]. Compleat [McR51]. Complete [Ano46-100, Mos91]. Complex [BES02, CHP71, Cae73, Dou69, Eas00, Fel83a, Mel71, Tsi72b, Vof72, AH91d, Cha88c, CN91, Den14, FK12, Gus04, KM89, OW90a, RE19, Sim11, Tak11, Rod70, Ano91-63]. complexes [Per92a]. Complexities [Wol71a, Wol71b]. compliance [Duf87, KH13, Jas81]. complicate [Gro86]. complicated [Pal16]. complicating [Sam22]. Comprehensible [Sno68a]. Comprehensive [Ano90n, Arz61, Fel75b, Zer13, DM81b, Eps86, Joh96b, Pom79, SL87, Ano79e, Ano88e, Kas21, Pif24, Wie82b]. Compromise [Lew64a, Sig88a, DS47b, Fie90, Ros78b]. Compromised [Ano46-30].

Compton [Dan57, Lan62, Tur54]. compulsory [Cal24]. Computer [Ano86q, Cha88a, Cha73b, Gol73a, Sha90, Wei78g, Sil71, Sil71]. Computing [OSS84, Ano98o, CO85, Kal85b, Mat97e, Orn85]. Comrade [Ano92u]. Comrades [But04, Day76a]. Con [DRM^{+}09, Hub60, Moo00b]. CONAES [Bro80a]. Conant [Gol95a, Wil71, Ano47p, Ber93b, Gol95a, Haw47, Wil53b]. concealed [DM24b]. Concept [And54, Ano48-32, Bor58a, Gee66, Sei73, Ano98o, GT14]. Conception [Mul65]. Concepts [Lev72a, Sag86, Smi61]. Conceptual [Whi71]. Concern [Alb75, Ein51, Tra73, Ano84j, Ano98l]. Concerned [CFG^{+}81, Dow83d, Sie72]. Concerning [Ano57k, Ano60d, Lap59b, Sha54b]. Concerns [Gle94, Kin83, Nai81]. Concerto [Zag93a]. Concessionist [Swo86]. Concessions [Sha73]. Conciliation [Hil65b, Alb83]. Concluded [Fed59a]. Conclusion [Ano60a, Int92, Isa91f]. Conclusions [Ano58e, Kel79, RGK95, Rot96g]. Concorde [G.84a, Ros78b]. Concrete [Ano63s]. Concurring [Mur54a]. Condition [Mel75, Rab68d]. Conditional [EP86, Tar23]. Conditions [Ano81c, Kow77, Rie60, Sab72, Cas77g]. Condliffe [Whe84b]. Condon [All48, Ano48n, Ano48w, Ano55f, Ano55i, Con48, Con55, HUC48, Hyn69, Pub55]. conducted [Alb98c]. conducting [Sme22]. Cone [Pri81b]. confederation [Isa94d]. Confer [Jac64b]. Conferees [Rus78]. Conference [Ach10, Ano45d, Ano45f, Ano45q, Ano46-46, Ano46-53, Ano46-59, Ano46-86, Ano46-90, Ano48-35, Ano49k, Ano50l, Ano50-45, Ano55l, Ano55p, Ano55q, Ano58k, Ano59d, Ano59i, Ano61i, Ano64p, Ano64q, Ano70a, Ano84d, Ano85f, Bev56, Cal61a, Car49a, Dra73, E^{+}84, Eps90b, Isa90f, Jac55, Jac63b, Jac65, Jac66, Jac73a, Jac71b, Joh10, MS77c, Nat86b, Ney73, Noy47, Rab58f, Rab62b, Rab67h, Rab70c, Rus56a, Shi56a, Top56, Ure45, Wei55a, Wei69a, Wei56b, Zei54, Ano58l, Bar80c, Bar80k, Cle89, Col88, Cor78, Goo85b, La 79, LR79, Pug80, Tov89, A.62, Ano57i, Ano58g, Ano60d, Ano63o, Ano72a, Ano79n, Ano82e, Bro58, CF74, Cla66, Coc59, Com57a, Com57b, Com57c, Eps75b, Fed59b, Fed59c, Hoc57, Int67, Jac58, Jac61, Ken79, Lal67e, Mah59b, Mah59a, Nor56]. Conference [Ore60b, Rab57g, Rab70c, Rap11, SW56, Top58b, Wei58, Wei55c, Mor79]. Conference/1974 [CF74]. Conferences [Ehr72, Rab60e, Mar70, Rab61k, Sim11]. Confidence [Ano54-28, Cal66, Lew07, Sho77, Bor82, Fis91, HK21, Len11, Wie84, Zuc89, dTdLF24]. confidence-building [HK21, Len11, dTdLF24]. Confident [Pic10]. Confidentially [HC87]. confinement [Don82, GG11a, Win81b]. Confirm [Ark85c, Han90a, Rob89]. Confirmation [Ano47v]. confirmed [DeW88]. Conflagration [Bla88]. Conflict [Ano49v, Ano84d, Ano90m, Ano10a, Ber78a, Ber77b, Bie68, Dam61, Dar51, Gla78, Hor68, Kre88a, Mye87a, All15b, All18, BO76, Cho86b, Dav77a, Dun83, Fai97, GN10, HSV19, Jon90b, Saa84, She61, Dam61, Sch63c, Was62]. Conflicts [Ede89, FKM87, Sub84, Sut84, Bus85]. Conformity [Hut53a, Ken53]. Confound [Wor69]. Confront [dF79f, Kri92b].

Confrontation [Bal67, Gal86, Rau99, Hol85a, Smi95b, Tro20, Van22, Wil82c, Bir87, Deu83a]. Confronting [Bul09d, Mer90]. confronts [CR17]. confused [Abl85]. Confusion [Bon49, Vil75, BRB16, Lov89, Pos85, Rot81a]. congenital [Ano10b]. Congestion [Kra73]. conglomerate [Mor85b]. Congo [Fin65, Pom66]. Congratulations [Cas85b, Gre84b]. Congress [DW85, Ein48a, Her88b, Isa97b, Lew68b, Rab62a, Ada83, Ano47z, Ano48i, Ano50-43, Ano55e, Ano87-29, Ano91n, Ano91-69, Ano97m, BWS^{+}80, Cas77a, Cha69b, Cla63, Cof67, Dad67, Din78, Fel76c, Fla78, Gol79, GG63, Hol86c, Isa83b, Isa84b, Isa85a, Isa86b, Isa86f, Isa87a, Isa87c, Isa90z, Isa91c, Isa91d, Isa92e, Isa97a, Isa97c, Isa97d, Isa99a, Isa02b, Isa02f, Isa03b, Isa03c, Isa03d, Isa03f, Isa03g, Jac62a, Joh94c, Lal64f, Lal66a, Lal67c, Low65, Mac86a, Mar49g, New62b, Pai83g, Pai85a, Pri56, Pro85, Rab49a, Ros89b, Smi75a, Spe87b, Uni78, Voh71c, Voh71d, Was97, Whi53, Dow80g, Got64, Kat57a, Div79]. Congressional [Ano48u, Ano49-36, Ano78g, RDS^{+}52, Ano45e, Ano46-31, Ano46-32, Ano48-36, Ano54h, Avi83, But99, Cas78b, Hol58, Ing60a, Kal60, Kal61, Nov56, Sco84b, Tru46, Wei78f, vN57]. Congressman [Mec23b, Heb51, Lap58]. Conjecture [Rit67]. connected [Ano93r]. Connecticut [Ano87n]. Connecting [AS06, Sto19d]. Connection [AH04a, Ben01a, Kah07, Sch75e, Wei85b, Ano83c, BP84, Bru90, EG24, FH83, FvHA86, Hol83b, IPB84, Kre87e, LLS83, PB84, PR14, Roc83b, Rog88a, Sal80b, Smi87b, Spi83, WPB84, Ano78n, Ber75a]. connections [Gle91, Nel81, MD81]. Connell [Ada87]. Connie [Dow81e]. Conquer [Mus07]. Conquest [Mor50a]. Conscience [Ano73d, Ben56, BBH^{+}73, Com51, Col06, Fel73d, Gro56a, Lap73b, Mei73, Sch66, dF79d, LL90, Pop81, Wil70b, Bre65, Tho62, dF79g]. Consciences [Sab76]. conscripts [Cra83d, Tsy83]. Consensual [Ste08]. Consensus [Bac77b, Nay12, Swe86, Abr79, AC85, RK88, Ros81]. Consent [Mav75]. Consequence [Llu99]. Consequences [Ano49-30, CP93b, Dow81h, Dow83b, Dys69b, Dys69c, E^{+}84, Fel76a, Fin84, Har81, Lea70, Lep62, Moo92f, Pat62, Tug51, You68c, dJ63, Abr80, Bac12, Coa47, Dro22d, Hue12, Mor12, RTB^{+}19, Sch98, vH11, Kri85, Rep85a]. Conservation [Ano78o, Car49a, Dix65, Ehr82a, Fre71, Gol61b, Hir73, Kra73, Pat79c, Rab69c, Sch80a, EE81a, HS97, Hir81, Joh81b, Pat78a, RW80, Sig87b, Sin76]. Conservative [Ing68a, Sto17a, ZS15, Orn67]. Conserved [Fri59]. Conserving [Fes77]. Consider [San76b, Bul01a, Sco08b]. Consideration [May47a]. Considerations [Ano53n, Bec64, Ben56, Cru67, Kal53, Mel53b, Ros60, KM18]. Considers [Lal64f, Mas48b]. Consistent [Kaz61, Ano93r]. consolidated [RL14]. Consortium [TV71, Hei11]. Conspiracy [Str05a, DS15, dF79c]. Conspiratorial [Shi54c]. conspirators [Ark92]. Constancy [Mes84b]. Constantine [RK93]. Constituent [May47b]. Constitution [Ano78m, Bor48, Gen84b, Mul55a, San68, Tel48a, Hen59, Loe87].

Constitutional [Dou24]. Constraining [Cra84b]. Construct [Ano49-41]. constructive [Isa88f, Bru50]. Consultant [Foz62a]. Consultants [Eid66]. Consultative [Buc82]. Consumer [Ban22, Cra82b, Eps77c, Gol82c]. Consumer-friendly [Ban22]. Consumption [Day74c, Pat82a]. Consumption-Control [Day74c]. cont [Ano96k]. contact [Nee79]. contain [KR12]. Contained [Sti83, Eng78]. containers [NA92e]. Containing [Rub08, Sto65a, You66a]. Containment [BvH82, Bou00, Kor10, Lal66d, She83b, Zub00, Gil18, Ken87b, Tur87]. Contaminant [Kin80c]. contaminants [MM11]. Contaminated [Ano69h]. Contamination [Ano91-61, Ber70a, Ito14, LT99a, Oga58, Ram70, LT99b, Tho80c]. Contaminations [Abb84a]. Contemplates [Que70a]. Contemporary [Cum98, Gen84a, McL66, Opp48b, Chi17, Lif13, MS77b]. Contending [Soc20, Iri86]. Content [Rab63i]. Contest [Rot01d, Bul02f, Bul04b]. Contested [Ber93b]. Context [Fis83, Rei01, Ano86y, Vog13]. Continent [Eml66, GD85, Ano85i, Fit11]. Continental [Ano54v, Ano67i, KH54, Tur53, WB06, Gué70]. contingency [Abr83]. Continue [And60, Jac64c, Mul52, Sez20]. Continued [CK20, Ano94m, Ano94n, Ano96j, Dic77, DR92, FB89, Moo91e, Tra71]. Continues [Ano67a, Isa04f, PWW63, Sto66b, Ark95d, Fri21, Gor96e, Har84b, WK82]. Continuing [Ano70a, Ark06, Bal60a, Bur90c, Cha86, Ste78, Bar77b, Lif13, Ano74i, Dal92b]. Continuities [Pla64a]. contra [NA87f]. Contraception [Con81a]. Contract [Ano49e, BN67, Bos51, Eri95, Nie66a, Rot79a]. Contracting [Yud83]. Contraction [Ste95]. Contractors [Ano89e, CLD88, Kub88, McG90, Ano87n]. Contracts [Har84a, Rei63, Fri79, HFMS79, O’S86]. Contradiction [Bro84b]. contradictions [Zim81a]. Contradictory [Was63]. Contrasts [Dav58, Ran85]. Contribute [Saw63]. contributing [She14]. Contribution [Mor57b, Mor79, Pri75d, Sac57, BW18, Nor56]. contributions [DM24a]. Contributors [Bul08g, Bul08h, Bul08i, Bul08j, Bul08k, Day77l]. Control [And71, Ano45a, Ano46m, Ano46-45, Ano46-106, Ano46-60, Ano46-57, Ano46-78, Ano46-72, Ano46-93, Ano46-95, Ano47d, Ano47-27, Ano48n, Ano48-40, Ano48-37, Ano49q, Ano61i, Ano65g, Ano66l, Ano67h, Ano71f, Ano85c, Ano86k, Ano89c, Ano89d, Arn47, Art70, Aue78, Bal89b, Bar46a, Bec71, BAB^{+}46, Ber89b, Bla69a, Ble89, Bor69, Bot61, Bow89, Bra47, Bre59, Bun89, Cou48, CLR74, Cav50, Cav56, Che88, Cho58, Cla74, Cle67, Cof70, Com57b, DB52a, DB52b, DB52c, DS47b, Dan72, Dav69, Day74c, DeW62, DR71, Deu67, Div79, Dje72a, Dje72b, Dje74, Don66, Dre89a, DuB69a, Fed58, Fed63, Fel61, Fel70b, Fel70a, Fel75a, Fis46, Fox46, Fre84a, Fre72b, Fry65, Gar89, GF73, Gom61a, Gor06, Gor67b, Gra89, Gra84b, Gre62, Gre68a, Gro46, Gro47c, Gro48, Hal89, Ham61b, Han46]. Control [Hen71a, Her85, Hig48a, Hoa64, Hoa72, Hod55, Ing52b, Ing57a, Ing60a, Isa91a, Isa96c, Ise92, Jac74b, Joh05, KFM50, Kin80b, Kin80a, Kin86, Kre89a, Kre93, Lal64d, Lal65f, Lal66a, Lal66c, Lap61a, LeR50b, Ler49a, Lev89, LW61, Loe84, Lon85b, Lut89, Man89, MT55, MF10, McD74b, McM48b, Mic61, Mor84a, Mot47, Myr72, Nel99, Nit89, Nye89, Oli48, Opp46, Opp48a, Opp63b, Opp63c, Osb47, Osb49b, Rab47k, Rab48e, Rab61a, Rab72b, Raj10, Ram72, Ras63, Rat71b, Ren99, Rod67, Spe58, Sci46, Sch89a, Sch68c, Sch63d, Sch71b, Sch79b, Shi48b, Shi48c, Sin05b, Slo68, Smi84b, Smi85c, Smi89, Smy60, Soh61, Spo72, Szi50a, Tay56, Tee65, Ure47, Ure48a, Van89, Wal69, Wal89a, War89, Wes74a, Wie70, Wil67b, Wil73b, Win84b, Wri65, LLR^{+}47, You65b, You66c, You68a]. Control [You71, Yut69b, Zlo47, dF79a, dN71, vdB65, vM89, AH20, APR17, Ano47k, Ano82c, Ano86s, Ano91z, Ano97n, Ano15a, Arb16, Arb19, AM86b, Ark84b, Ark87a, Bar90b, Bar85c, Bel20b, BPG89, Bor82, Bre13, BMPd89, Bug22, Cha87, Cla81, Dan16, Dea87, Dje81, Dre82a, Eis90, Fei92, FP17, For91, Gad89, Gar84a, Ger13, GK21, Gor84, GK85, Hal83, Han96a, Har84a, Ibr13, Isa81a, Isa85a, Isa85d, Isa86c, Isa87a, Isa87b, Isa87c, Isa88b, Isa88c, Isa92j, Isa95e, Jac83, Kar88b, Kin78b, Kin84a, KB89, Kis79a, Kor19, Kre86b, Kre87e, Kre88d, Kre90, Küh17, Lam87, Lee91b, Lei78, LST10, Lin84c, Loc93c, Mac88, Mac86c, Mat85, Mec20b, Mec20g, Men86, Men96b, Moo14, Mor87b, NA91o, NA92a, Oel12, Pai85e, Pai85c, Pan86b, Pat80e, Rab50e, RR88, RKK22]. control [Sch77a, SD86, Sea89a, Sha81b, She85d, She86b, Tro20, Tuc10b, Vad21, War78, You78a, Zer13, Ano46-63, Ano60f, Ano87n, BGR61, Cir07a, RK89a, Ano86k, Ano89g, BGR61, Cav61, Div79, Kru74, Lan56b, Lep89, Pri70, Tho80g, Hac81]. Controlled [Fed59b, Kur56, Ler49a, Lil46, Tel56, Ano84p, Lil62]. controllers [Ano90r, Tsi89c]. Controlling [Abb84a, Bla84, Leg56, Rot77, SSK10, Yor73, Uni78, Fra83, She79]. Controls [Ano45p, Ano46m, Ano69f, Ano88e, Bec55, Bel57, Cav52, Fis46, Gaf65, Ing59a, McC48, Mil50, Alb94a, BB20, Cha85b, CC13, Nov77b]. Controversial [Gil83, Rot79a, Cas78a]. Controversies [dF80]. Controversy [Ano53-47, Ano72g, Kea61, Sen47, Nel74b, Pet61, Pol57b, Rab56a, Rum71b, Sag69, Sin81, Sto61, TG70, Tam71c, BHM91, Cam79, FB89, Gre78, Hig77, Hol76a, Mot83, RK87b, Sch78b, Ste90b, Tho80h, dF80, Sch68e]. conundrum [Gol95a]. convenes [Cor78]. Convenient [Sch99]. Convention [Ano57j, BST02, Ben01b, BHH11, Cas74, FCC^{+}46, Gol78c, Len11, McL11, Sim11, Wri46, Fra75, Gol83a, Int78, Isa96d, JH16, Wal55, Ano72b, CK20, Mar64c, Mar64d, Rab46a]. Conventional [Ark85a, Asp87, Ber84a, Boy10, Cot84, Dea89, FLLW91a, Gol07, Møl90, Pol09b, Sen87, Sig87a, Smi93d, Ano87-33, Ano88h, Fer10, For85, KB88, Mea85, MH89a, Rob78c, You87a, Alf84, Cro65b]. Convergence [Lov78a, Tuc10b]. Convergent [Lew71f, LM71]. conversation [Alf77, Ano12g, Jr18]. Conversations [Ano63k, Hei70, Rab63e, Hei71]. Conversing [Got65]. Conversion [Alb93f, Ano93w, Ano94j, Ano94m, Ano94n, Ano96k, Ano96j, Har90a, Isa93c, Lal66b, Pal01, Pre92, Wei64a, Ada86a, Ano97l, Bea86a, Bea87, DG86, Khr94, Uhl96, Wol76, Kel79, Kli81, Uhl96]. Convert [Jac74b, Nol17]. Converting [Bev86, Day77c, TBBS85]. Convictions [Pol50b]. convocation [Ano83h, Kal82e]. Cook [Eri95]. Cookbook [Dan05, Dan20]. Cooke [Tic94]. cookie [Coh87]. Cooking [Com75f, Moo18]. cooks [Kre93]. Cool [Oel20, Spa96, Dro25c]. cooled [Ram16]. Cooling [Jac74a, SS88, Zac82, Hir76]. Cooper [Bei84, Gom71]. Cooperation [Ano46-27, Ano50-49, Ano51l, Ano63p, Ano64o, Ano10a, Bak60, Ben01d, BM08, Cha60, Cra64, Cre63, Dad71, Eme60, Foz62h, Fru63, Hai62, Hek60, Ken63a, Kow55, Lal66c, Lal66e, Lev70, Lov63, Mal50, Mar70, Mar68, Nac65, Rab46e, Rab56e, Rab64f, Rab65h, Rab65d, Sab63, Sch62c, Sch63e, Sea63b, Smi69b, Smi70c, Sto72, Tho53, Tod64a, Top62, Wad58, Wad63, Whi64, vdH61, BV10, BHK^{+}21, GN10, Hum66, Jam86, KSG15, Mat85, Rap11, Dew55]. Cooperative [Dou60, Neu65, RC70, Ric67, JFS20, Smi16]. Coopersmith [Dro24d]. Coordinating [Pow47]. Coordination [Foz62f]. Cop [Moo01b]. Copenhagen [Pet81]. Copernican [Sha54b]. Copernicus [Bou74, Cha73a]. Coping [Mic73, Mye92]. Coppock [Wil81a]. Cops [Moo03b]. Cordiale [AB05]. core [Rad97]. Corinne [Ber77b]. Corn [Nel77]. Cornell [Bal79c, Ano46-33]. corners [Rot96h]. Corning [Ano55i]. cornucopia [Pat79c]. cornucopian [Bar78a]. Corollaries [Ber89b]. Corona [Mat97c]. coronavirus [Ger20]. Corporate [Har89a, Mil69b, Sch75e, Smi13, Dou15, KH13]. Corporate-funded [Har89a]. Corporation [Mil69a]. Corps [Ben61, F.61, Yar80a, Ano62-28, Sch69b]. Correct [Gol88c]. Correction [Ano46-34, Ano50m, Ano50n, Ano52f, Ano56f, Ano61h, Ano74c, Ano75d, Ano80h, Ano89n, Ano90-52, Ano92-30, Ano92-31, Ano93t, Ano93-30, Ano94p, Ano94q, Ano94r, Ano95p, Ano97q, Ano19a, Ano19b, Ano19c, Ano20, Ano22a, Ano22b, Day77d, Gou75a, Lev90, NA89d, Rab56d, vH72, vH76a, PW76b]. Corrections [Ano78j, Ano90-54, Ano90-53, Ano93-31, Ano94s]. correlates [SS24]. Correspondence [Bea47a, Ras83, Ano46-29]. Correspondent [DeB81a, Osn77]. Corrigendum [Ano12c, Ano14a, Ano17b, Ano18]. corrosion [PR14]. Cosmic [Ber75a, Fla90]. Cosmogony [Bur65]. Cosmology [Ano74e, Ste74]. Cosmonauts [Chu94b]. Cosmos [Fla90, Rau69]. Cost [Ber70a, BF73, Ble89, Dys75, Han90c, Har70, Kno66, Lew70b, LC09, Mat57, McD08, New73, Ram70, Sim60a, Uhl98a, Ano87j, Ano11h, Bos98, LN86, Mat12]. Cost-Effectiveness [Kno66]. Costa [Ano91u]. Coster [Nor04]. Coster-Mullen [Nor04]. Costing [Isa01e]. Costly [Day93, Isa93f, Lum96, Gol86b, Mec21f]. Costs [Ano46-74, Bin04, Gri65, Isa03a, NA89i, Sch47, Ano88i, DM81a, HCH16, Sch98]. COSWA [Ano62u, Ano62x, Rab63e]. Cottrell [Nor64]. Could [AS02, Ano78o, Arb91b, Fri48a, Moo15, Rob75b, Wol80, Alb06, Che52, Hil12, Isa87e, LP10, Man77, Pos21, RTB^{+}19, Roc76, RH17, Rot94i, Wu22]. Coulson [Sto58]. Council [Ano51i, Ano80f, Ano89f, Gol88g, Har79a, See79, Ano77l, vL92, Ano46-35, Ano46-41, Ano47f, Ano47j, Ano47s, Ano85g, Got50, Got51, Osb58, Pai82j, Szi62b, Zim75]. Counsel [Bur50, Hym91]. Counsels [Rod85]. Count [Cum99b, Jac74b, Ano87-31, Ano90-52, Bar77b, Has85, Nye89, PD89, Squ12, WH89]. Countdown [AH04b, Bar98, HC12, Jon13, Opp02a, BM22, Bri86, Dow82h, Dow83d, Dow79b, Lac85]. counted [Bar81e, Eps94]. Counter [Fry56c, Kre96, Van87]. Counter-Establishment [Van87]. counter-revolution [Kre96]. counterforce [Oel12, Tsi82c]. Countering [NA87f, Tsi82c]. Countermeasures [Ham63, YSW21]. counteroffensive [KB88]. Counterpoint [BR90, FB89, PD89]. counterproductive [Wol80]. counterproliferation [Gol94]. counterrevolution [Isa91i]. counterterrorism [Ano94o]. Countervailing [Gan84]. counterweight [Moo95f]. Counties [Cro65a]. Counting [Ano95i, Dea04, LC98a, NK15, RD77, Ano90a, Nuc95, NA88a, Nye89, Per88b]. countless [HBD17]. Countries [Ano90-55, Bau60, Bro71, Coh72, Ded57, Deu60, GS09, Ikl60, Kap66, Kih47f, Mah59b, Mah59a, Myr71, Pre60, Ric67, Sim46a, YES^{+}21, Zah69, Cli77, Kap76, Kib12, Nay12, Squ12, Tud21, Ano78o, Tel76]. Country [Ano60f, Ber72a, Gil72, Gol66, Ing59b, Mot47, NAB99a, PK63, Sma76a, Coo78, Gar78a, Not15, Roc18, Ash81, Bro55b]. County [Rei63, Sin79a]. Coup [Ahm00b, Jac92]. couple [Fla90, Isa88c]. coupon [Ano98o]. Coups [Hun87]. Course [Ano83i, Cir07b, Cum01b, Mul68, Tho80c, Ure47, Whi64, Ano83r, Joh81b, Mal81]. Courses [Lau64]. Court [Ano96-31, Kan77a, Ano90-40, Day77m, Kan77b, Kri78, Men86, Sen47, Moo96n, Ano87n]. Courting [Bro89b, Gla76b]. Courts [Rab53c, AB77, Ano87y, NP80]. Cousteau [Dan90a, Isa90d]. Cover [Fre95, Rot05a, Sha90, Sin91, Ben79b, Hun85b, Mat97c, Mei90]. cover-up [Hun85b, Mei90]. coverage [Ame86, Fri11]. coveralls [Ano89-27]. Covering [Ano14b, Hal93a]. Covers [Nor88]. Covert [Chu76, Lei92, Mec16b]. coverup [Ano85o, OH85, And78a]. COVID [FWCL20, Mec20d]. COVID-19 [FWCL20, Mec20d]. cow [Ano91-27]. Cowardice [But99]. Cowboys [Mol03]. Cowering [Pen77b]. Cowpies [Dow83c]. Cows [Wil75a]. Cox [Day77b, Ano88e, Mes83b]. cozy [Cas77a]. crack’d [Rot99a]. crackdown [Ano90-42]. CRADAs [Rot94b]. Craft [McN59]. Crammer [Sha50]. Crash [Cra76]. Craters [Szi71, Wes71]. Cravens [Ano15c]. Crawford [Ano88d]. Crawling [Rot94b]. craze [Mec19e]. crazy [Mat98e, Sto23]. created [Nea80, Wag61]. Creating [Sai10, Sha22, Soc15]. Creation [Bur65, Pla59b, Sch07c]. Creative [Cra61, Rot01b, Etz77, Mat97e]. Creativity [Eas85, Hol59, Oro59, Ten75]. creator [Eav17c]. Creators [Gol88a, Pee78]. Credibility [Ack87c, Sau75]. credible [SF81b]. Credit [Cam95a, Gle52, Rab47a, Rab51i, Ano87-34, Car03, Flo89a, Isa92m]. Credo [Gou63, PMRP50]. Credulity [vHWD76]. Creek [Ano89j, Ber73b]. creep [Ham93]. Cresson [Smi83]. Creveld [Ste89e]. Cries [Pea88]. Crime [Ano93s, Str06c, Cre68]. Crimes [Har88c, Joh97a, Woo01d, Iwa82, Jac82a]. Criminal [SO06]. Crises [Osg67, Pom65, Dea91, Leb85]. Crisis [Ano71c, Ber75c, Bin67, Bla84, Bor70, Bro67b, Cor58, Cre74, CP62, Das96, Dav72, EE86, Ell71, For51a, For51b, Fre76a, Gla62b, Gla68, Gol70b, Gri84, HG84, Har62, Jac74b, Kas15, Lam83, Leb91a, Leb94, Lew07, Man59, Mar61b, Mor65, Mos61, Nad98a, Ols74a, Pal64b, Rab46g, Rab59a, Rub74, Spi71a, Sta73, Ste76a, Szi60b, Tel66b, vH77a, Ble87, Blo85, Bro79d, Cho86a, EG86, Etz77, Fig23, Gor81, Hes84, IS15, Ken80, KB80, Lüb90, McK23b, Mot89, Nea80, Pie19, Red82, Wil85a, Wu22, Zuc89, Bla84, Lyn69, NK12a, Ano90o, Bla84, Leb92, Mil69a, Mye93b, You82, Bla82a, Pom66, Sti83]. CRISPR [BV20, OGFM18]. Criteria [Ano49b, Ano55a, Ano89c, Bal89b, Bar67c, Ber89b, Ble89, Bow89, Bun89, Coh62, Dre89a, Gar89, Gra89, Hal89, ILA70, Kre89a, Lev89, Lut89, Man89, Nit89, Nye89, Rab69g, Sch89a, Smi89, Van89, Wal89a, War89, vM89, Ano77n, Dah78]. Critic [O’B69a]. Critical [Ano74d, Bey85, Don86b, Her83a, Kra10, Lew70e, McN72, PvH72, She83b, Spe72, Isa85d, MD24]. Criticism [Jac62a, Lap59g, Wig60]. criticisms [Kan77a]. Criticized [MW47]. Critics [Sag69, Ack90c, Ano90-63, Pai83d, vH85]. Critique [Bro47a, Cam77, EH72a, Fru75, Kar80, Shi47d, Spo72, Fre81, Gom81a, LL83, MP78, Abr74]. Crop [MW81, Wes81, Wil77a]. Crops [Jac73b]. Cross [Ano86i, Bea67, Fra91, Ark95c, Rot97j, Ano87s, Con88]. Cross-Cultural [Ano86i, Bea67]. Cross-Examining [Fra91]. cross-hairs [Ark95c]. Crossed [Ano78p, Ark98a]. Crosses [CB06]. crossfire [Obe98]. crosshairs [Ano10c]. Crossing [Ano70c, JOO15, Wil75e, Eng91]. Crossroads [Ano10a, Joh70, Key80, Kis72, Mar48d, Moo94d, Pol69, GN10, Jac79, Ano46-71, SH94]. Crouch [Hol83a]. crowd [Ger13, Ibr13, Joh89, Zer13]. crowded [Sch77b]. Crowe [Ano14c]. Crowfoot [Ano90o]. crows [Moo94a]. Crowther [Fie19]. Cruise [Fie90, Gor06, Loc89b, NA90q, NK07b, Tsi75b, Tsi88, Ano88p, Ano88t, Ark89d, Fel88, Hun90, Ing84, Mac88, NA87e, Par17, Sha84e]. cruisers [NA90m]. Cruisin [NA88g]. Cruit [Smi83]. Crunch [Spu73, Zam96b, Sch95a]. Crusade [Ano78k, Gom71, Szi47b, Usb47, Dys81, Spe95a]. Crusades [Eav16]. Crusading [Dow79a]. Crushing [Mul48]. Crust [Vol74]. Cruz [Pat90b]. Cry [Spi54, Wil90]. Crying [Ehr72]. Crypto [Arn22]. cryptocurrency [Dro22d, VU22]. crystal [Rot99a]. Crystals [Smi54b]. CSCE [Luc90a]. CTB [Arn96a, BV97, Cal80, Eps93, Eps96a, Zam96b]. CTBT [Han05, HGV10, Kre99, Pod24]. CTR [OG10]. Cuba [AG63, DFF^{+}63, Gar88c, HB63, Ing63a, Leb88b, Leb92, Pea92b, Rab63b, Smi95a, Wri97, Leb94]. Cuban [Leb92, You76c, NK12a]. cuckoo [HJS^{+}92]. Cuff [Fel83e]. culpa [Ano96o]. culpable [Kri98]. Cult [Pat83, Sub87]. Cultural [Ano86i, Bea67, Ein52b, Ham61a, Noy46b, Pla64a, Whe84b, Ano69d, Wei83a, Ano69c, Ano69s, Cha69a, Git69, Jac74b]. Culture [Bar66a, Boy91, Boy05, Fal61a, For86, Gee66, Lak64, Man64, OW90a, Opp59a, Ray81, Sie72, Sim59b, Smi87c, Sut54, Tae51, Ano91-28, Azh93, Bro13, Smi87a, Tre21, You87b, ZA04, Hol83a, Els90, Mul73b, Pee78]. Cultures [Sim65]. Curate [Gil65, Sim65]. Curatorship [Kat95]. Curbing [Isa88f]. cure [Ano98n, Ste77c]. Curie [Rab58b, Ano49-27, Ano54p, CCD^{+}52, Rab50d, Rab54c]. curiosities [Ano94n]. Curious [Alb93f, Mar64g, NA91e]. Curiouser [Ano94l]. Current [Ano46-36, Ano53m, Ano53l, Ano54i, Ano68d, Ano68e, Ano68f, Bor68, Bre68, Bre69a, Cha68a, Com54, Cro68, Dev69, Doy53, Dys69a, Eis59, F.61, Fel68c, Hau68, Hib68, Kuz58, LB53, Law69, Lew68c, Lew69b, Lil53, McG68, Mur54b, Ney69, Rea68, RKR55, Ros68b, Rus55a, S.55b, Sch68a, Sha54a, Sim54b, Sim55, Sim59b, Sim59c, Sim59e, Sim59d, Sim60b, Sim60a, Smi58c, Smi58a, Smi58b, Stu69b, Uhl03b, Wie70, WF68, Wil68a, Wil68c, Wol68, Yat69, You68b, Coh86, GP84, Wei17, Fel72a, Rot62]. Currents [Sut68b, WF16, Sha90]. Curricula [Ano88e]. Curriculum [Gil72]. Curtailed [Ano53-45]. Curtailment [Ano50o]. Curtain [Ano51j, Boy87, Hus03, Pai83c, Shi52, Shi87, Har86a, PP83]. Curtis [Ing70, Lap69, Gro56b]. Curve [Fol74]. Curveball [Alv05]. Curves [Com75f]. Custer [Rot04d]. Custody [BM02, Smi90b]. customer [Ano95i, Fer10]. Cut [Ano50x, Gor95a, Rot91b, Wal06, Wri99a, Kau90, Men89, Smi21]. Cutoff [Zam95c, Lan89d]. Cuts [Ano64i, Ano97r, Ano88q, Flo89a, For85, Gar88d, Gay83, Isa94f, Kok88b, Loc91a, RK89b]. cutters [Coh87]. Cutting [Hir81, SG69c, AM21, BR90]. cutting-edge [AM21]. CW [Sim60c, Spi72c]. CWC [Bar10a]. cyber [Bre13, Car13, Gor96c, GCV16, Kor16, Lin19, Nye13, Seg13, Sme22]. cyber-enabled [Lin19]. cyberattack [Sha18]. Cyberattacks [RK19]. Cybernetics [Hoa64, Rap65a]. cybersecurity [Lin12, WC17]. cyberspace [Ano11f, CG12, CC13, Dem19]. Cycle [Bla61b, CL06, Hol74, AF88, Ano96n, CJLT24, Kaw78, Len11]. cycles [Dah78, DLF79, Lov79, vHWL79]. Cyclotron [DuB49, Mar54e]. Cyclotrons [Nis47]. Cynic [Lev63]. Cynical [Pri76b]. Cynthia [Roc85]. Cyprus [Bar98]. Cyril [Lal67a]. Cyrus [Her83a, Pol79a, Rab60e]. Czech [Wes03]. Czechoslovak [Ano90-47]. Czechoslovakia [Ano64q, HJ76, Jac69, Jan76c, Mul69]. Czechs [Hib95e]. Czeslaw [Fer54b].

D [Ano78k, Bro05, Bus85, Can70, Div79, Dys72, Fis83, For87, Fre84a, Gen86, Gou78b, Gra62, Haw56, Hec70, Kar80, dK60, Lan90k, Lap57, Moo95c, Nor06, Pat81a, PF75, Rab72a, Rub73, Smi83, Sto59b, Wel79, Win84b, dF79g, dF79d, Sco08b, BN67, Cob88, Dro24e, HV18, Hol64, Mar64f, Nel64]. D-5 [Cob88]. d-word [Sco08b]. d’être [Pug73]. Daalder [Bre87]. dacha [Sch89b]. daddy [Fly99c]. Dael [Ano60f]. Dag [Rab61c]. Dagomys [Pug88]. Dahlman [Dav79b]. Daiichi [vH11]. Dale [Lap69, LP92]. Daley [Gre73b]. Dallaire [Sch06c]. Dam [Hal75b]. Damage [Ano91z, Cog82, MNS46, Mul55b, Nel99, Ric99, RSW55, Blu23]. damaged [Fal89]. damages [SS90a]. Damaging [Pan07b]. Dammed [Mos67]. Damn [Uri94a]. damned [Ano98k]. dampens [Car84a, Tho84a]. Dams [Ada73]. Dan [Dow84b, Kre83a, Mos91]. dance [Hud78]. Dancers [Tho85b, Ano86j]. Danes [Bur87, Nat88]. Danger [Ano74i, Coh76b, Gov93, Hud70, JC50, Kri75, Lal64g, Lew61a, Lyf61, Oue07, Phy80, Phe61, Pot49, Rab53g, Rid47, Rot94c, Rus54, Sei49, Sto65b, TK66, APvH86, Ano93-34, Den82, Fel78j, SD10b, Sut16, Wat77, Wil92a, Wri91, Men96b, Lap69]. Dangerous [Ano46-37, Ano89o, Ano91q, Ano98k, AK98c, Ble87, CA98a, Cob88, Cul79, For06, Gla76a, Hol94a, Lei92, Nel00, SM05, Tel47c, Tel82, Ano83c, Bea87, DS15, Fel80j, Goe80, Hol83b, MvE13, Mec19e, Pos87, Sha18, Val97a, FH83, Lei92]. Dangerously [Can72, Ger72, Rab72g]. Dangers [Ano50p, Ano58l, Bus47, Fel72a, Lap59b, Mar87a, Ore58a, Rab59n, Roc81, Arb77, Gal76, Leb87, Pol80, Rop16, Smi16, WF16]. Daniel [Boy95, Bro55a, Bro60a, Fel77o, Gen84c, Mar79a, Men91d, Mye89, Pat83, Ros67, Saw79, Sha91b, Sno68b, Sti83, Tel49, Yar80a]. Daniels [Bro54, Ano46-39, Ost72]. Daniloff [Lar87, Sha86]. Danish [Dra01, Nie14]. Dante [BBG^{+}85]. dargestellt [Thi46]. Dark [Eli85, Gen86, Joh97a, Rho95, Smi85d, Ano92v, Cap77, E^{+}84, Han92b, Dow84b, Fel76i, Han95a]. Darkness [Aus99, Lap59h, Fri20]. Darlington [Rot98a]. Darmstadt [Som83b]. DARPA [All17, Ano90-50, Arn92b, Der85]. Darstellung [For80]. Darwin [Elm71, Gre54, Mul73a]. Data [Bru58, Cro58, Hum59a, Lap59a, Neu58a, NE76, SL58, Swe89b, You58, Ark87f, BBP14, Gof93a, Isa90e, OW90b]. Database [Ano87r, Ano88d]. Databook [Ano84g, Cal84a, Han94, CAH84, C^{+}87a, C^{+}87b, C^{+}89, NBF94]. Dates [Kir91]. dating [Coo79, Mar82b, Mat56]. Daughter [Blo05, Moo00a, Pal00, Mar83c]. Daughters [VR67]. David [And65, Ano85c, Ano88e, Ano91r, Bam85, Beu49, BGR61, Bou98b, Cog82, Col84b, Cra83a, Cra84d, Deu83b, Don86b, Dow81b, Dow81e, Fel83d, Fel83e, Fla90, Fre84a, Gan83c, Gre67b, HD82, Ham83, Hos87a, Hun87, Jac74c, Kre89c, LaF90a, Leb94, Mar78a, Mer79, Nei82a, Ols62a, Pat81a, Pee78, Pom67, Pos79, Rab68g, Ras63, Rec89, Rot93a, Rot98a, S.80, Sch68f, Sta85, Sti83, Stu89, Swe93, Tho80g, Tod79, Wal81a, Wal89b]. Davidson [Kat54d]. Davies [Ano78n]. Davis [Dow80c, Dro20a, Kin86, Szi50b, Wei55b, Wil69b, Wil80b]. Dawkins [She79]. Dawn [For86, Rab60c]. Day [Ack99, Ano87n, Ano89f, Bal79b, Coo75b, Day77b, Eng91, Mes85a, Ols59, Sch74b, Shi84, Wag61, Wea82, Aba89, Ano98k, Cla77, Fri20, Moo96h, Rot96b, Spa96, Eng91, Fel66b, Ham91c]. Days [And73, Flo88, Weg96a, Nea78a, Nea78b, Shi84]. De-alerting [vH97a]. De-netting [Ano94j]. De-politicizing [Bou98a]. Dead [Goo03, Jos61, Smi85c, Wie67, Loc89b, LC98a, Bri86]. Deadline [Kre99]. deadlines [Isa88a]. Deadlock [Fed50, Fri48a, Rab46c, Som83a, SS90a, Abt84, BW84, Car83b, Gay83, Hop84, Ing84, Kel84, Lod84, Lon84, Pei84, Saa84, TT83, Wie84, Yor83, ZG83]. Deadly [And78a, Dow81c, Goo03, NA90d, Tho79b, Ber87a, Arm68, Mei90, Sto85, Swe85]. Deal [AH01, Fly05a, IF51, Isa02a, Kri02a, Sig98, Cal97, Har90d, Isa89k, Ise97, Kib10, Kru81, NA93l, Pai85f, Sha95c, Shr23, Sig87b]. dealers [Ano90-46]. Dealing [Law88, Mec19b, OW99, Orl66, Sob91a, Pie19]. Deals [Kum04, Lum94b, Zar76, Ack90i, Fer10, Gol90a, Har87e, Men89].

Dean [Dow84d, RS52, Sha91b, Ach56, Ano52j, Bir87, Mur58, Sha95b, Smi54c]. Dear [Eat78, Fel79b, Gol21, Law84b, Rus78, Ada95, BWS^{+}80]. Death [Abb04b, Ber04a, Cal17, CF74, Ded58, Fre95, Gig60, Har92b, Isa91e, Lif69, Sha90, Sko67, VS04b, Woo55, Wou73, Alb98b, Ano94j, Bar77b, Bon57, Bon58a, Bon58b, Dav86a, Gol96a, Gol95d, GM90, Hud78, Leb94, Zag95a, vHJ97b, Dro22e, Hec22]. Deaths [Ist67]. Debacle [Pat79a]. Debatable [Isa96e]. Debate [Ano47g, Ano53m, Ano67a, Axe51k, But73, Cha75b, Coh74, DeW60, Fox46, Ful69, Gar87, Gom65, Ing79b, Isa89h, Lal67d, Mar75, Mar48b, Net75, Pal64a, PvH74, Rab60b, RA67, RK89a, Shi47c, Sim59a, You67b, dR49, vH75a, vH75b, Abr79, Ada83, AG86, Aga79, Ano96t, Arn80, Ber87b, Bey12a, Cha82c, Cha86, Cle11, CO14, Day77j, DR92, Dow77, Fai80, Fei77, Fox88, GP84, GP92b, Gou77, Gra81b, GH83, Har79a, HH87, Isa89i, Isa91f, Isa92f, Isa94g, Jan15, JL87, Kri78, Lan90j, Ler81, Lop78, Mar89d, Mye86, NK86b, Par87, RR15, RDS85, RAB^{+}80, Sch13a, SF81a, Sch79c, SS76, SD87, Tsi87, Zha15a, vH90a, vH91, Ano81f, Cra83b, Rea80, TD89, Tho80g, Fel81b, Dow81g, LaF62, Mil85c, You65a, Kal68]. Debated [Ano49-36, Hig48a, Isa96e]. Debates [Ano47v, Ano48j, Ano50q, Ano53k, Lal67c, Shi48b, Val86, LLR^{+}47, AJSB79, CD80]. Debating [Ano89-28]. Debever [For80]. Debit [Rab47a]. Deborah [Ano86l, Mea93]. Debris [BB10, Pri02]. Debt [Ano93m, Fly01b, Gra76b, Kra91]. debunk [Mec21a]. Debunking [Gur88b, Uhl98b]. Decade [Ano47n, Bre93, Cav56, Gro56a, Leg56, Mar96, Rab60c, Rab70d, Smi70c, Sto72, Isa88f, Küh17, Ano57c]. Decades [DeB81a, Pot70, Ano87x, Ano13d, Fel82a, Jan15, Mil86a, RR15, Shu87, Str62, Zha15a]. decarbonize [Pie16]. Decay [Alv00, Ein52b, Rei93]. Deceit [Mei90]. Deceive [Ric03b]. December [Ano61i, Chu54c, Eis54c, NCA88a, Tel57a, Ano66s, Lea97, NA88e, NA89h, NA90e, NA91f, NA92g, NA93h, Str62, Wat82]. Decent [Bra58]. Deception [SvH77, Wel80, Alb98d, Ber86a, Hof86, Mit97, Han92b]. Deceptive [Gus01]. Decide [Com50, Pas77, Wei68a]. Decision [Ano46-61, Ben01b, Ber93b, De 74, Dow81a, Dow83a, Fel67b, Fer62, Fri70a, Gee71, Gom67, Hig66, Hol66, Ing56b, Isa00e, Jav68, Kre90, Mar59a, Opp50, Pau52a, Rab47i, Sar85, Scl82, Smi58d, Sti47, SZC54, Str05b, Wal90a, Wei71, Ano80r, Bak86, Bro23, Fed76, Fel79c, Fel81j, Hol76a, Hou85, Kis76a, Mac98, Mes85b, Moo97m, Ros78b, Ano78n, Gib82, Pat81a, Ber77b, Dow82h]. decision-maker [Fed76]. decision-makers [Kis76a]. Decision-Making [De 74, Scl82, Hol76a, Hou85, Ano78n, Gib82, Pat81a]. Decisionmaking [Gil68, Mor62c, Kre85a]. Decisions [AP80, Ano49p, Bro67a, CH71, Har62, Wol68, Isa86a, Squ17b, dF80, Smi83]. deck [Joh95, Smi94a]. decked [Knu89]. Declaration [Ano52g, Bok49, BBF^{+}57, GG87, Pon82, Uni78, AGC^{+}88, Ano58k, Pug88, Ano58l, Ano72h, Ano75c]. Declarations [Fra91]. declare [Sim84e]. Declared [Ano54b]. Declassification [Ano46-38, Ano46-75, Ano50p, Hut46b, Jon13]. Declassified [Ano46-51, Ano46-52, Ano47m, Bru90, GP92b, Gre47, NK16, Han90b]. Decline [How65b, HK13, Kun97, Ros81, Squ13, Wil87]. Decommissioning [Ano94l, Loc14, PR85, Woo01a]. Deconstructed [Col00]. Deconstructing [Suz11]. DeConto [Dro24c]. decontrol [Mus80]. Decoupling [Bre72]. dedicated [DM24a]. Deed [Ale67, Ano97i]. Deeds [Ade72, Sta60a, Ano96o]. Deep [Dow81f, DH90, Hal75b, Mat98c, McK74, Ano92v, Gay83, Han93, Ano97r]. Deep-six [DH90]. Deepening [For51a, For51b]. Deeply [Ano47x, Kyl97]. Default [Ano59a]. defeat [Isa91o, Nel92b, Ple96]. Defeatism [Coo58, Ple96]. Defence [Ano87n, Bre69b, Sim85, Som84b]. Defend [Ano46o, Ano53-40, Mei53, Rob90b, Sin79a]. Defended [Ano53-45]. Defenders [Rab60d]. Defending [Lib01]. Defends [Pal76, Sch75b]. Defense [Adv60a, Ass63, Ack97, Ada52, Ada87, Ano48n, Ano48-32, Ano50i, Ano50j, Ano50k, Ano50h, Ano50-44, Ano51e, Ano51f, Ano51g, Ano51k, Ano51n, Ano53n, Ano54v, Ano54x, Ano62o, Ano62p, Ano78o, Ano86j, Ano97l, Ano10a, Ant50, Ark01f, Aug48b, BW84, Ber84a, Bet69b, BCMR92, BBZ01, Bon49, Bon50, Bon51, Bou61, Bra49, Bre67, Bre72, Bre73, Cha53, Cle01, Cof65a, Cro64, Cro65b, Dea52, Dee59, Dol76, Dow81d, Dys64, Dys69a, Exe57, Ede89, Eva04, Fed62, Fel57, Foz61d, Fre84a, Gai69, Gro48, Hal76b, Har88d, Haw51, Hir88, Hoh75, Hop49, Hum55, HUC48, Hut50, Hut51, Ing54a, Ing67, Isa82d, Isa92f, Isa92g, Isa98b, Isa99b, Isa99d, Isa00a, Isa01b, Isa02c, Isa04c, Isa04f, Jaf51a, Jaf51b, Jan50, JL74, Joh01, Ken71, KH54, Kna63, Kno55a, Kre87f, Kup51, Lal64f]. Defense [Lal66b, Lal67g, Lap50, Lap54d, Lap58, Lap59c, Lap59d, Lap60b, Lar50, Lev52, Lev53, Lew61a, LS65, LP08, Mac62, Mar65f, Mar66a, MFC62, Mar67b, Mau50, McM65, McM51, Men99, Mil85c, Mor62d, Mue01, Nas82, New62b, Osg62, Pal01, Par53, Per20, Pie62, Pit64, Pla53, Pod01, Pos00, Pot53a, Pot53b, Pug05, Rab50a, Rab51g, Rab66c, Rat58, Rav65, Rei63, Rin57, RK89a, Rot93f, Roz64, San57, Sau01, Sch80e, Sel00, Sha65, She00, Sim56, Sim60c, Siv80, Sle52, Sle53a, Sle53b, Sle53c, Sle53d, Sle53e, Sle53f, Sle53g, Sle54a, Sle54b, Sle54c, Sle54d, Smi53, Spi72c, Bul01b, Ste85a, Sto53, Stu89, Sub68, Swe86, Sym50, Tos82, Uhl96, Ura01, Was62, Wat53, WB85a, Whi76b, Wil68c, Win84a, Wit03a, Wol82, Wol63, Wri63a, You67b, You68c, Yud83, Zei50a, Zei50b]. Defense [Zei51, Zei52, Zei54, Kos04b, Ada83, Ald76, Ani15, Ano86p, Ano87w, Ano88o, Ano91-42, Ano92-30, Ano18, Arb18, Ark86i, Bin18, Bos81, Bru88, Cot84, Daa87, Dea88, Eps87a, Fel77k, Fla88, FC84, Gap90, GN10, Gho11, GP18, Gov93, Gra84a, Gre83a, Gre18, GU88, Gro84a, Gur88c, Her88c, Hou85, Isa81b, Isa84a, Isa85c, Isa94a, Isa98c, Kap78b, Kap78c, Kap83a, Kap89, Kar17, KWM88, Kor21, Kre91b, Kre92a, Lam87, Lea87, LP10, LvH18, Man79, Mec18b, Mec21c, Mec21d, Pai82i, Par87, Raj14, Ras80, RWK^{+}83, Riq15, Sch19, Sen87, Sha87a, SD83, Sim87b, Smi84c, Sok87, Som83a, Sør81, Spa95d, Stu87, Thr15, Wig66, Abl85, Ano50-29, Ano87i, Ano93y, Ano93q, Ark95d, Dow81g, Dow82h, Dow84c, McN67a, McN67b, Mec19c, Mil86b, Nor64, Rep86, She60, Wri63a, Kol84]. Defense [Asp79, Dys72, Rep86, She82, Sig85b, Tay87, Wol62, Yon68, Sco79, Sel63, Som84b, Uhl97a]. defense-or [Ani15, Riq15, Thr15]. Defense-protected [BW84, WB85a]. Defenses [For02, Lac69, Moo02c, Rot05b, BV10, BS97, KK19a, Ada87, Ano87m]. Defensive [Abt84, Cla54, Sei53, Hol82, Hop84, Zam95b]. Defensor [NA93j]. Deferment [Axe51a]. Deferred [Zei64]. Defiance [Sme76]. Deficiencies [Pri47]. Deficit [AW85]. Define [Eme47]. Defines [Ano47x, Hor60]. Defining [Ano87o, Hai81, Nei88, Sch01a, Tho85a]. Definition [Ist67, LM75, Mur55a, Win71, Lal77, Rot95c, Sta81]. definitions [Dav83]. Defuse [Mie74]. Defusing [Ram04, Ric02]. Degradation [Col74, D’A70]. Dehousing [Moo95h]. Deibert [Ano11f]. Deitchman [Cro65b]. Deitrick [Gol92a]. Déjà [Ano94k, Wei04]. Del [Day88, Ano91u]. Delay [Fed59a, Kih47f, Osb47, Smi76, The84]. Delayed [Ano46-57, NBM85, Joh83c, Moo96c, Zam95c]. Delbert [Pla57b]. Delegate [Gro47b, Ano94l]. Delegation [Ano60e, Ano60c, Ano60d, BAB^{+}46]. deleted [Isa90e]. deletion [Sch96]. Delhi [GG87]. deliberate [Ark85a, Ekl77]. deliberations [Ano47f]. Delicate [She07, Voy76]. Delight [Sam06]. Delivery [Dot65, Lal65g, Nix10, Mor84c, PN09]. delphic [Dro23k]. Deltas [VSD^{+}09]. Deluded [Bal60b]. Delusion [Bal60b, Gus83, Mar59b, Pai05, Sto77, Ark85a]. Delusions [Khr94]. Demac [Mar85a]. demagogues [Kea24]. Demand [Bas80, Bat61, Chi59a, Har88b, HGF17, Lan89f, Sez20]. demanded [SN89]. Demanding [Ano97l]. Demands [Ash48, DS47a, Mas48b, Ure48c, Wie48, vL48b, LaF90b]. Demarcation [Men96a]. demining [Bou98a]. Demise [Ano82j, Sno81]. Demobilizes [Mor46]. democracies [Ano91-44]. Democracy [Ack90d, Ano47x, Boe52, Dei05, Dou48, Ing56b, Mar75, Nad66, Nas04b, Uhl04a, And94, Cha89c, Kea24, Man84c, MD24, Nad98a, Per92a, Sca88, Scl82, Wal83a, Lyn69, BW83, Gib82, Eld92]. Democratic [Ano59c, Aur71g, LS92, Moo51, Smi67, FP17, Gou24, Sha22, Mar64c, Rab48b]. Democrats [Isa03b, Zag93a, Zag95e, Ano60n, Ano84l, Mag88b]. Demographic [Rea80]. Demographics [Cho86b]. demonology [Jac82b]. demonstrated [Pal77a]. Demonstrations [Mil69c, Sel69c]. Demonstrators [Mil70]. Demophora [VT73]. Dems [Isa04d]. Demystification [Moo99c]. Denaturing [ABB^{+}46]. Denial [Ung78b, Ano10c, Eva54, Jac15, Per13, Sic96a, Wea11, Moo95e]. Denied [Cos02, Rot97c, Ano94n, Har96]. Denis [Ano78m, Bri86]. Denmark [BØ79, Nat81, Sør80, SS80]. Dennis [Cha74, Tsi89a, Dro24a]. dense [Pai83g]. Denuclearization [Bar67b, dF79b]. denuclearized [Han90a]. Denuclearizing [Hec08]. Denver [Ack99, Ano45f]. deny [Ark85c, Kri92b]. Department [Adv50, BE62, Lan90c, Rot79a, Sti83, AK93, Ano46-54, Ano47-28, Ano87i, Ano88n, Ano93y, Ano93q, Ark87b, AC65, Bro47b, Bro23, Dre51, Dro16a, Eli46, Hoe87, Ing57a, Kil65, Mor55, Mor85b, Nel93b, NA91g, Ram88a, Tod64b, Wee98]. Dependence [Sti78, Ano86p]. dependency [Isa88f]. depends [Dav83, Kun23]. Depleted [Ark99e, Mes97, Die91]. depleted-uranium [Die91]. Deploy [Fel67b, The84]. Deployment [Men71, ASB^{+}84, BR85, Dre87, Joh83c]. Deployments [NAB99a, NK09b, KN14f, KN17e]. Depolarizing [Gro78c]. deposit [Fre77b, NA90l]. Depository [ZSA73d]. Deposits [Bra50]. Depression [Rei68a, Isa88c]. Dep’t [ABB^{+}46, Ano46-81, Ano46-89, Ano46-88, Ano46-95, CCMU46, Shi46b, Tel46b]. Dept. [Ure46c]. depth [Dan90a]. Deputies [Ano91-40]. Deputy [Fis61b, Ham91a]. derailment [Sch81b]. Derek [Opp62a, PF75]. Derian [Haf79]. desalination [Rab16, Ale69]. describe [Mec19c]. Describes [Nis47]. Descriptions [Abr04]. Descriptive [Ada61]. Desert [AH01, Hod69, Sha69, Ano93-27, Ark93b, NA92a]. Desertification [Ano97l]. deserts [EB78]. deserved [Fer54a]. deserves [Fed76, KT97]. Design [Ams56, Mul66, Ric53, Rot89a, Bul02f, Ano97l, DM85, Tel47b]. Designation [Pan07b]. Designed [Ano88e]. Designer [All18, NA90k, Hop77]. Designing [Mel00, Nor11]. designs [NA87g, RLG11, Wri87]. Desist [Bet95]. Desktop [Ano98k]. Despair [Dro24e, Mar96]. Desperately [Ric00a]. Despite [Ano74i, Bug22, Eps75c, Fly98c, Isa90t]. destabilization [Gal78a]. destabilizing [Cob88, Say93]. Destiny [Soc15]. Destroy [Pan92, Kea24]. Destroyed [Cah76, NA89j]. Destroying [San14]. destruct [Cle89]. Destruction [Ano78n, Com51, Cap98, Coo81, Ehr81, Gla76b, Nis47, Nor64, Sta71, Smi05, Wel60, Yor74, Alf79, Hoa77, MW81, PC92a, Rob78c, RT12, Smi93b, Tay96, Wes81]. Destructive [FD76, Ric75b, Dor96]. Details [Mar97a, Nor04]. detant [Fra76a]. d’État [Hun87]. detect [Roc76, Tur18]. Detection [All59a, Ano58e, AIL^{+}62, Bro60c, Bro60d, Ham58, Mar94, Ore58b, Ric98a, Wig58, FJ12, Jeh60b]. Detective [LM71, Zei73]. detectors [Sch24]. Detente [Ano74i, Eps75c, Jac64a, Lal64e, Gar77b, Haa77, Hor88c, Mie79, Osn77, Nat88, Sto78a, Sto79, Woo99a, Gal86, Pat79a]. deter [LJ17, Pan22, San14, Wu23]. Deterer [All65]. Detergents [Way65]. Determinator [Sie05b]. determine [Bar77b]. Determined [Dem19, Mec20e]. Determinism [vB67]. Deterrence [Ams56, Ano85b, Bar67c, Bos81, Cof65b, DFJ^{+}63, DJE^{+}82, Dub85, Jer77, Kel62, Kri02c, Küh18, Lew08, Lin82, Mac61, Mus99a, Mus76, Nit89, NH08a, Per85, Rab57e, Rab59a, Rat58, Sco74, Sin61, Sny85, Ste60b, Tay87, Yor74, You76a, AC02, APR17, Ano82j, Ark86i, Car83b, Doy19, Eps77d, Fel78b, Hol16, Ing85, Jer85, KB88, Kor10, Lal77, MS85, Mcc86c, Mea85, Mil84, NH08b, Sco08b, Sno81, Tal18, Tar23, TY75, Väy87, vH90a, Alf84, Arm68, Han90c, Mas66, Ras63]. Deterrent [Aro60, Bro63b, Buc60, Kno60, Mur72, Ros75, Sle56, Szi64, Cot19, Tur13]. Deterring [All65, DE23]. Detlev [Han76]. Detonating [Bur05]. detonation [Bor14, Hau17]. Detonations [Ano60r, Bro63a]. detonators [Wio81]. Detractors [Gol62c]. Detroit [Day75b]. Deutch [Gra94]. Deutsch [Ano81c]. deutschen [vL47, vL48a]. devastating [FD23]. Developed [Tel76, Cli77]. Developing [Bar61a, Bro71, Coh72, Dje68, Dow67, Dun66, Gil72, Jon70, Kap66, Lev72a, Lon69b, Myr71, Pre60, Ric67, RE68, Sal64b, Was66, Lop78, Mar54d]. Development [All55a, All56l, Ano49u, Ano55g, Ano60f, Ano60k, Ano66d, Ano76b, Aur71c, Bac48, Bac49, Bad66, Bau60, Ber73b, Bro58, Bro67b, Car69, Coc50, DZA47, Dea53b, Dys60, Fra66, Fuc70, Gou75b, GS09, Gre64b, Gro53b, Heb52, IW49, IW52, Jan64, Kam79, Kno66, Lap59a, Lau47, Lop66, Lop68, Mab69, Mah59a, McK74, Mei54, Mei61, Mil50, Mor79, Opp47b, Pal64b, Pla64b, Pre62, Ram64, Sal64a, Scu65, Shi61, SIF66, Smy50b, Sol66, Spo51, Sto72, Sui71, Tel76, WW61, Zuc52, dF79g, Ano98p, Ark84e, Bar78a, Gor81, Lak82, Lan85, Mak76, Mar83b, Mei77b, MM86, NA87h, NA88q, Pug76, Per78, Tho82, Tsi77b, Tsi83d, Wio79a, YG77a, dC78, Ano50u, Ano52k, Ano78n, MS77c, Pla57b, dF79e, Kon67, Mun54]. Developmental [dRVT^{+}47]. Developments [Ano54u, Ano55s, Ano61s, Ano62-29, Fel72a, Foz61d, Foz61j, Hin53, Kow48, LB53, Lea70, Ran55a, Sim54e, Sta60b, Wig51, NA88c, Raj14]. Device [Gom74, Old62]. Devices [Tri77, Dro22f]. Devil [Ano90-35]. Devine [Ing79b]. DeVolpi [Gre82b, Zim81b]. Devolution [Hun87]. Dew [VS04b, Cap02]. Dewey [Ano48r, Ano48-30, McM48a, Tru48a]. DeWitt [Kat57b, Ano56c, Ano89p]. Dexter [And55, Mag81]. DIA [Kre87b]. Diagnosis [Rui68]. Dial [Moo97c, Mec21h]. Dialectical [Hol53]. Dialogue [Ano56m, Com72, EH72a, Man80, Sch65a, Sha54b, Eva75, Mar87b]. Dialogue1 [Hsi65]. Diamat [Eps52]. Diamond [Cow65]. diamonds [Ano91-35, Ust93c]. Diane [Dow80e, Mec21c]. Diaries [For51a, For51b, Joh51]. Diary [Ano48k, Oli66, Szi50b, Hor77, Nea78a, Nea78b, Bru55a]. Dibb [Jor87]. Dibblin [Ham91c]. Dick [Wil52]. Dickson [Moo02d, Col84b, Gor73]. Dictionary [Ano87s]. Did [Ano90-31, Ano91-36, Ano93z, Bor65b, Coh93b, Err58, Gar88c, Gol98, STS^{+}49, Wal89c, Ack89c, Alb89c, Alb89d, Ano88v, Ano91-39, Ano95t, Bea86b, Cla90, Dro23b, Eve88b, Fly99c, Gar98, Isa94c, NA00d, NAB00, AK98a]. diddlers [Gof93a]. Didn’t [Ano99-30, Wei94a, Ano91-31, Ano95l, Bro88b, Flo88, GW98, Kat90b, MHM90]. Diebold [Sil71]. Dierkes [Wil81a]. Dies [Ano78n]. Dieter [RK87a]. Difference [Haa73, Isa91a, Isa02b, Kra72, Isa88f]. differences [GG11b]. Different [Fel92, Fer88, Gan76, Smi85b, Ano84i, Jr22c, Mel78, Møl90, Zal19]. difficult [Mec21e, Roc76]. diffuse [Kis76b]. dig [Wei91a]. Digging [AH92a, Cae74]. Digital [Tsi75a]. Dignity [dF79d]. Dijkstra [Dro25b]. Dilemma [Bou87, Cha69b, Con65, Cra84e, Dra01, Fin54a, Fry56b, GA74, Hil52, Hor66, Hym69a, Ing71, Kes85a, Lew63b, McA65, McC48, Mus76, Sau78, Sen63, Ste95, Szi62b, WI71, Wel60, Ano81j, BB22, Fel77n, Fel78b, KKL15, Kre81, Lef77a, Mec18b, Par78b, Rem15, Sch15a, Sha87d, WP80, Dow80g, Gai92, Gan83b]. Dilemmas [Pol09a, Rea80, Sco74, Wil68a]. Dilks [Bam85]. Dillon [Fly98a]. Dimension [Bam85, Deu83b, Sno69]. Dimensional [HE72, Lew72b]. Dimensions [Sto56b, Ano12b, Lif13, Wei76b, Sti83]. Diminish [McV71]. diminished [Lon95]. Diminishing [Ada06, Lal64g]. Dinner [HUC48, Ano98n, Moo18]. Dino [Leb92]. dinosaurs [Man23]. Dioxide [Doc07, FS81, JP13, Sto19a, WMRK79]. Dioxin [Ano80h, DD79, Tho80c, Tho79b]. dip [Mat98d]. Diplomacy [Bad66, Bra89, Cra61, Har69a, HBM^{+}07, Hig66, Mar49a, Sin63b, SNB^{+}54, Tod78b, War63, Alp85, Bar10b, Bro86i, Got20, Kri92a, NA88c, Pic10, SW21, Wei86b, Cra83b, Ham61a, Kol63, Sti52]. Diplomat [Gya90, Men92b]. Diplomatic [Mak06, Mar65e, Som05, GP18]. Diplomats [Hur60, Men69]. Direct [Ano78n, GG11a]. directed [Dow79e, Dow79d, Dow79a, Dow80a, Dow81e, Dow82d, Dow82g, Dow83b, Dow83e, Dow84d, HD82, Hec19, JFS20, Wea82]. direction [Moo97e, Phe56]. Directions [AK98c, Fit66, GB73, Ing82, Kne65, Sup67, Cas78a, Rom89]. Directive [Mem80]. Directives [Ano50-47]. Director [Ano49-32, Asp79, Eav17b, Gol88b, Ekl77, Han98a, Mec23a, Ano49u, Ano50-48, Ano52k, Mec21c]. director-general [Ekl77]. Directors [Ano89v]. Directory [Ano87n, Ano89g, H.84b, Sal53, Ano89f]. dirt [WM78]. Dirty [Har88c, HLL03, Koc04, WS91, Woo00f, Zel76b, Gea89, Han89, Pai04a]. dis [Ano98p]. Disadvantages [O’S63]. Disagree [GA08, Ste75b, Ano90-63]. Disagreement [Cav56]. disagrees [Ano91-32, DT91]. Disappearance [Har81]. disappeared [Lau91a]. disappears [Eve89]. disappointing [Kap90b]. Disappointment [Dha90, Har59]. Disarm [AHK90]. Disarmament [Adv60a, Ano47f, Ano47g, Ano49-42, Ano50-39, Ano50-45, Ano54u, Ano55q, Ano55y, Ano60j, Ano61i, Ano62q, Ano63-28, Ano66l, Ano71b, Ano72b, Ano78n, Ano78m, Ano81i, Ano86l, AH62, AE85, Arz61, Bec71, Bet62b, Bet70, Boy10, Bur62a, Cav52, Cla63, CY81, Dre07, Dul57, Dys63b, Ehr82c, Eme63, Fel61, Fin57a, Fis62b, Fis62c, Fre09, Fry56c, Fry56b, Fry56d, Fry57b, Fry57a, G.84b, Gla62a, Gol72a, GG63, Hin18, Hum58, Ikl60, Ing52c, Ing54b, Ing55b, Ing63a, Jac62b, Jac63a, Jac64a, Jac65, Jac66, Jac73a, Ken60b, Kih47f, Kin80b, Kin61, KH63, LaF62, Lal64g, Lal64h, Lal64i, Lal64b, Lal64c, Lal64a, Lal65d, Lal86, Lap61a, LS65, Loc94, Loe84, Lon64, Mar65g, Mar66b, May08, Mye90, Nea63, Nix60a, NA91a, NA91c, NA92j, Ore63, Pug76, Pat62, Pie58, Pie60, Pom61, Pri58a, Rab57a]. Disarmament [Rab61a, RV64, Rob75b, Rus62a, Sen57, Sar78, Sch59a, Sco08a, Sco08b, Sen65a, Sha60, Sha61, Sha81a, She64a, Sim54d, Sim59a, Slo68, Smi84b, Soc55, Soh61, Soh62, Som84c, Spi58, Ste77b, Ste52, Sto64, SG62b, Szi55a, Tal65, Top58b, Van46, Whi66a, Whi66b, Win94, Win98, You63a, You66c, Zam95c, Ach10, Al 15, AGC^{+}88, Ano75g, Ano76d, Ano87-28, Ano90-60, Ano96l, Bar76d, Bin15, Can79, Cor78, DS47b, DPT17, ER84, Fah90, Far77, For91, Fra76a, Gra84a, Hen96, HGF17, Her88a, Her79, Isa91g, Jac77, KR11, Küh17, LS92, Lei78, Lun79, MW98, Moo97o, Mos89, Ple96, RR81b, Sak79, Sel21, SD15, Smi84c, Sto79, Tho83, Tho82, TM79, Ulg15, WH11, Ano52r, Ano78l, Ano78m, Bar82a, Fow55, Ham58, Hud78, Sha78a, Sha78b, Wei78a, Ano85d, Bra89, Cav58]. Disarmament [Dav69, Ing82, McG84, Ras63, Sco74]. Disarmed [RP66, Fel78c]. Disarming [Fin65, NA93f, Pai85a, SB23, Kre83a]. disarmingly [NA93f]. Disarms [Wur94a]. disarray [Ack87g]. Disaster [Ano70e, Ber73b, Bro89b, Fel79d, Gid73, Got87, Joh00a, Mic73, Par80, Rei78, WKW57, Wyl57, FK12, Hen17, Ito14, Kle91, Rot93h, Hos87b, Rei78, Hos87a, Nor64, Wal81a]. disasters [Ack89h, AH89b, Møl90, Pat90b, Tin23, Ano55c]. disastrous [FK12]. Disavowing [Tho80i]. discharge [Ano94n]. Discipline [Bro68, Ano94m]. Disclaimer [Nov76]. disclosed [DeW79]. Disclosure [Lev60, BJCP88, NCA88b]. Discontent [Alv01a, Kib01]. discontents [Küh18]. Discontinuance [Ano59i, Ano60d]. Discord [Isa89a]. Discourse [Ber77b]. Discourses [Gus01]. discovered [Mer24]. Discoverers [Gen84c]. Discovery [Eas85, Kih47f, Lew73b, Fla90]. discriminate [KB88]. Discuss [Hig50, HM72, Ran85, Say84]. Discussed [Ano50-46, Ano46-56]. Discusses [WC47, Leg87, Mec21c]. Discussion [Ano50t, Fox46, Gra84b, WZSA73, WK82]. Discussions [Jac64a, Sch47, Shi48c]. Disease [All68, Gol95c, Lwo60, Cro18, Ano78k]. Diseases [Baw60, Kap60, Kap66, Sal63a, Hum77]. Disenchantment [Sin63b]. Disengagement [Bre72, Osg60, vW59, Lod84]. Disengaging [Kal85a]. disguised [Smi97]. Disgusting [Cam53]. disinformation [Ano11c, Fie21a, Mec21a, Mec21h, YWS20, YSW21]. Disintegrating [Ber93a]. disinvention [Mac96]. Disjunctive [Bar66a]. Dislike [Hor60]. dismal [Pre96]. Dismantle [Sen87, Ber23]. dismantlement [Alv14, Hic22]. dismantles [WM78]. Dismantling [Bir87, NK04a, Sut91, Gol93a]. Dismay [Lew76]. dismiss [Mec21a]. Disorder [Ano93l, Woh64]. Disparities [Hus04]. Dispel [You87a]. Dispelling [Ack87a, FS81]. Dispersal [AM51, Ano54j, Aug48b, Aug48a, Aug50, Bol51, Gor51b, Haw51, MTK46, May52, MM51, Mon54, Tru51, Whe51]. Dispersion [Ano51p, Ano51v, Gor51a, Lap51a, You65a, Zim55]. Displacement [Osg62, McL18]. Display [Bai75a, Bai20]. Disposal [AK01, BKSG60, Gao17, Hun84, MvHKN01, Maw77, dL59, EG24, HC71, Kan11, Law84a, Moo97m, Swe79a, TPPB21, Tra24]. disposes [Gar83b, SG83]. Dispute [Bey80a, Jac73b, Mac65a, S.80, GP18, Rub87, Swe78b]. disputed [Gel90]. Disputes [Mou73, Wei69b]. Disquiet [Buk97a]. dissemination [Fra47]. Dissent [Day75c, Gre73c, Sci92]. Dissenting [Hym68, Ros69, Smy54, SW56]. Disservice [Ano54c]. dissident [Ano92u, Ham91a, Mar78b]. dissonance [DH98]. Distance [Sut54]. Distant [Tho80d, Moo91b, Rot95m]. distinctions [Phi93]. Distinguished [Ano95j]. Distorted [Mus99b]. distorts [Ber82]. distraction [Squ21a]. Distress [Wag83, Bud22]. Distressing [Bud22]. distributed [Dow83a, Dow83f, Dow83d, Dow84b]. Distributing [NR79]. Distribution [Ano47i, Ano47-29, Ano48z, Ano49y, Gre47, Lib58, NAD^{+}47, Rob49, Uni47]. district [Ano03i, Isa94f]. Disturbing [Ste76c, HSV19]. ditch [Wil92c]. Ditched [Cap02]. dive [ZC92]. Diversification [Pan23]. diversion [vH85]. diversity [Fel77h, Sto18a]. Divide [Haq05, Mus07, WB06, Woo93, DM18, Ikl80]. Divided [Isa91f, Ker06, Sim59b, Sul58a, Jac82d, Sim90a, Tan83]. dividend [Ack90k, Ano94l, Ano95i, Moo95o]. Dividing [Les93a, Mur55b]. Divine [Chr71]. Division [Ano52k, McK51, Pfe51, Sal53, Pos21, Ano49u, Ano49-32]. divisive [Bro89a]. Divvying [Isa91g]. Dixon [Ano12f]. Djerassi [Con81a, Smi97]. DMZ [Woo98b]. DNA [Lef77b, Tho80g, WSH^{+}84, Bal78, CT78, Day77g, Elw64, Fed78, Gre78, Kri78, Nov77a, Sal77, SSK10, Tha78, Tho79c, Vic78, Wri77, WS83]. Do [Ano49h, Ano58l, Ano65h, Ano90o, AM74, Bar57, Bra13b, Cal61a, Cha63a, Cir00, Com75b, Fel83d, Fis62a, GLSS07, Hen93, Isa04d, Kat61b, Kre95, Kur79, Lil49d, Lon67, Lon73, LC93, Lor01a, Mac11, Mei48, Moo92g, Noy57, Orm57, PMRP50, Ros90b, Rub74, SOE21, Smi83, Web06, vW58, Ada82, Alb89c, Alb89d, Ano87k, Ano91-69, AB18, Coo78, Dro22g, Dro23i, Fly99a, Fly99c, Gar93, GW98, HSV19, Int81, JS79b, Kap80, Lam83, Lee24, LSP95b, McK23a, Mec23b, PC92b, RE19, Row77, Ste89d, SD10b, Woo92, YES^{+}21, Zha19, Zwe84, vM89, Ano93u, Pat80b, Ano91-40, Bar80l]. Do-It-Yourself [Cal61a, Lor01a, Ros90b, Fly99a]. Doak [Iri79]. Dobzhansky [Cro67, Har64b]. dock [NA93i]. Docking [Rau69]. Doctors [Jos81, Zhe90c, Lor77]. Doctrine [Cra83b, Rab60a, Sin63a, SM05, Gar88b, Kin78b, Zys17, Rab47l]. doctrines [Roc81]. Documentary [For80]. Documents [Ano46-51, Ano46-52, Ano47m, Gre47, Pop52, McK74]. Doder [Tum87]. dodgers [Isa89k]. dodges [Ano91-54]. DOE [Ben78]. Does [Ano54c, Dan52, Dar71, Isa88e, Kol64, Ler49a, Mac67, Man66b, McV71, Sch59b, SMRVS22, Ten69, Alb94d, Ano91-42, Ano92v, Ano93u, Arn91, BW18, Eva75, Isa89f, Kib12, Nay12, Sco90, Squ12, vH16]. doesn’t [Hil16, Hil20]. Dog [CO74]. Doggle [Cow65]. Dogma [Wil53c, Wri49, Mcc86c]. dogmas [Isa89l]. dogs [NA90c]. Doin [Mat98d]. Doing [Ano94j, Dro24a, Pat78a]. Dokumente [For80]. Dolan [Ano78f]. Dollar [Ros75, Ark96e, CP93a, Isa93a, Pre92]. Dollars [Ano94t, Hol80b, Ano91-42, Lon86a, Nuc95, Sca88, SD10b, Woo98a, Ano96o, Rep86]. Dolphins [Pag61, Hut61]. Domain [Ano98k, Kal68, Mat97b]. Domenici [Cul08, Dom98]. Domes [Her78a]. Domestic [Ano46-36, Ano46-61, Ano46-96, Cow65, DB52a, DB52b, DB52c, Fis83, McC48, Tel47d, Uni47]. Domesticate [Rev66]. Dominance [Mar98c, All15b]. dominate [Dem19]. Dominates [Ray81]. Dominators [Kre03]. Dominic [Dow81f, Sto19a]. Dominion [Rot02g, Uhl04a, Ano98k]. Dominique [Hal93a]. dominos [Sha93b]. Don [Cum98, Gen84b, Hur78a, Cha78b]. Donald [All59l, Ano81f, Ano86i, Cha58b, Jas81, Klo85, Lap61a, Pit91, Ros68a, Sch69d, Sco74, Smi85d, Sta64, dF79f, dF79e]. Donaldson [Fis83]. Dónde [Fly02a]. Done [Arb91b, Asi57, Dow80e, Ise97, Rot04f, Cla92, Cla20, Fre91b, Pie22]. Donella [Cha74]. Dong [Lyn78]. Donna [Mar85a]. Dons [Bus46]. Don’t [Ano91-27, Ano94k, Ano95i, Ano95j, Ano96m, Ano98l, Ard93c, CG04a, Chu05, Coo76a, Cum99b, Ear86, GM87, Gar92, Gar97c, GF73, Has85, Kar97, Kre91b, Lyt85, Mil03, Moo96b, Mur54c, Nye89, Pat80b, Pod05, Sai98, Str90, Sut89, Tov89, WH89, Zel76a, Zha05, And93, Ano90-43, Ano91-36, Ano98p, Dal22, Dun90, Isa82d, Lat91, Rot92i, Rot92p, RC88, ZS15, Moo15]. doohickey [Ano90-35]. Dooley [Doo60]. Doom [PS72, Sch72c, HF80]. Doomed [Kum04]. Doomsday [Boy89b, RAS04, Smi83, Spa96, Sue82, Cla77, DP12, Gei16b, Mec19c, Sut16, Day77b, Kra10, SS24, vH89]. Doomsdays [Bul04a]. Door [Ano64n, Myr75, Rab62c, Dan91, Fel79d]. Doors [Str06a]. Dope [Har88c]. dorm [Ano92s]. Dormant [Kre86a]. Dornan [Jac80b]. Dornberger [Fis55]. Dorothy [H.84a, Nor88, Sim84b, Wal83b, Wil74b, dF80]. Dosage [Tam74]. Dosages [NW70]. dose [FR97a, FR97b, Hil12, Tho12, Wil92a]. dosimeter [Sch89b]. Dots [AS06, Ano93w]. dotted [Ark98a]. Dotto [Ano86i]. Double [Ash76, CGO05, Isa91h, Pea92b, Rab53g, Rab68a, Tho80g, Wal87b, Wei80b]. Double-Edged [CGO05, Wei80b]. Doubleday [Tel49]. doubletalk [Gar86]. Doubt [Vis56, Dow80i]. Doubting [Kwo10]. Doubts [Ano53-41]. Doudna [Mec20f]. Dougherty [Swe86, You68a]. Douglas [Dow79d, Hal93a, Hol83a, Mil85b]. Doumani [McK74]. Dount [Ste78]. Dove [Isa89k, Isa88e, Ano89j]. Dover [Pra04c]. Doves [Fel75b, Isa92h, Wal85a]. Dowling [H.84b, Hos87a]. Down [Cob78, Com75h, Hun03, Pai04b, Sch03, Sma77, vH97b, BW84, Bro93, Deu84b, Eps92b, Haf84, HI94, Ial18, Isa89m, Isa91m, Isa98b, Joh94c, Kor17, Loc92a, Loc95, Lyn84, McG90, NA89j, NA91b, NA93l, Pai83f, Pra98, Rot93l, Rot93u, Sch76a, Sha80, Ste90a, Ste85a, Swe79b, Swe87b, WB85a, WM78, NA89o, Bet89a]. Downed [Pri02]. Downing [Tsi83a, Joh01, Dre84b]. downloading [Mar91b]. Downs [Cla74, Mar91b, NA89s]. downside [Lac17]. downsize [Den14]. Downtown [Wat73a]. downwinders [Ano91-46]. Dr [Ano05b, Dow82c, Dow83b, Sch80c, Goo04a]. Dr. [All48, Ano48n, Ano48x, Ano50r, Ano52y, Ano57f, Con48, Con55, Exe54, Goo04b, Lib57, MJ68, Mar48e, Mor48a, Opp51b, Szi47c, Szi50b, Ure48b, VFIS48]. Draft [Bur90a, Com48a, FCC^{+}46, Nat68b, Rab46a, Tel48a, Wri46, Ano91-54, Isa89k, Mil83b, Por19]. draft-dodgers [Isa89k]. Dragon [Alv87, Kus83, Rab98, Sal78a, Mar58a, Whi89]. Dragons [Gif13, All17, NA87f]. Drain [Ded70, Mor71, Mur70, Cav90, Ung85]. Drake [Sha54b]. drama [Gus11]. Dramatization [RD75]. Draw [Bro87a]. drawing [Cho95, Zel92]. Drawn [Joh81a, Tay21]. Drax [Koe22]. Drax-tic [Koe22]. dread [Eri91]. Dream [Elk62, Gre63a, Gui77, Lan83a, Rob91b, Ano13f, Eis84, Lei83b, Spa95d]. Dreamland [Rot98a]. Dreams [Jos02, Rus62b, Wea63, vH77a, Hen17, Lon95, NA88n]. Drell [Fre84a]. dress [Dav92]. Drew [Sie07b]. Drexler [Bul07d]. Dribbling [Loc93a]. Drift [BPR52, Day75h, Gla76a, Ark85h]. Drilling [LH70]. Drip [Rot93g]. Drive [Ano91-40, Bal86a, DeW84, Eav18, Gho15a, Hoa77]. Driven [Fri54]. driver [HK13]. Drivers [Mas57]. Drone [She05a, Cal17, Gei16b, Sut16]. drones [Kun23, Lac17]. Drop [Dow82h, Hig66, AH92f]. Dropped [MHD^{+}05, Wea76, Ano91-33]. Dropping [Moo98e, San76a]. droppings [Gov93]. drops [Alb93e, Alb98a]. drought [Oba85]. Drug [O’B69b, LP92].

Drugs [Fre68a, LP92, Rad92b, Wil99, Ark86g]. drum [Moo91b]. Drumming [Rot92e]. Drums [Cap02]. Dry [JvH09, NA93i, RL14]. Dry-Cask [JvH09]. Dual [Ber04a, Bor70, Hec06, Hoa57, Hoe09, MF10, Men53, SG69b, Len21]. Dual-Use [Hoe09, MF10, Len21]. Duarte [Bul09b]. Dubious [Ano03g, Kra07, vCM99a, Ble89]. Dubos [Tod81, Rus62b, Hul76]. Dubrovnik [Ano63o, Woo96c]. Ducking [Isa94g, Ano91y]. duckling [Spa95a]. Dud [Isa02a, Kri02a, Hib95a]. Dude [Rot02c]. Dudley [Ano87-32, Mun54]. Due [Rab51i, Cam95a, Coh76d, McL18, NA89i, Wei93, Gle52]. Dueling [Isa89b]. dues [Sch92b]. Duffield [Joh51]. Duffy [Win84b]. Dukas [For80]. duke [Rot95d, Sha92a]. dumas [Zag94d]. Dumb [Ark00g, Ano95i, Moo92b]. Dump [Kap90c, Rot92m, Ack94, Ano90-27]. dumped [Zan91]. Dumping [Abb84a, Tay91, Ano91-58, Les93d]. dumps [Blo95, Jag95]. dumpsites [Ano90-41]. Duncan [Kin80b, Kre85a, Gre71]. Dungeons [NA87f]. Dunn [Fra83]. DuPont [Ano49n]. Dupre [Lew62a]. Durable [Nef78]. duration [DL21, Twi23]. Durch [Lac85]. During [Ano57c, Pol68, NK16, RK89a, Rub08]. Durwood [Dro19]. Dusko [Tum87]. Dust [FvH99, Wil86a, Cha89a, Zag92]. Dustup [Mak06]. Dutch [NR79, Ran53]. Duties [DuB48]. Duty [Ano53-46, Lev51a, Vis56, Isa91h, NA89j]. Dwight [Nor64]. Dworkin [Ano85b]. Dyad [Ric09]. Dycus [Ack97]. dying [Les93c, Meh95]. Dynamic [Gol67b, Hay76, PF75, Ben76b]. Dynamics [Jav68, Mec23c]. dynamism [Alb87a]. Dynamite [Don90]. Dyson [Ano93n, Dys65, Gom84a, Rom89, Tay69, You69]. dystopia [Mec21h]. Dziak [Cra83b].

E-Mail [Fer03]. E. [Bea47a]. Each [Gra81a, Ano98n]. Eager [Wer56]. Eagles [NA91g, NA92k]. Earle [NA87h]. Earlier [Ano86r, Per92b]. Early [Agn82b, And73, Ber93b, Boy86b, Div86, Far86, For86, Jos87, Wei94d, Wit01a, Ano94j, Bro79c, CYP14, Dre87, Dro23h, Sic96c, Vil75, Wit01c, YWS20, Tel49]. Earth [Ano86i, Bal79b, EE85, Hol86b, Lan58, Par73, Tod81, Woh78, Gle94, Isa91m, Pet96, Alb61, Ano55h, Ano82e, Ano14b, Bal79c, Ben72, Dra73, Far98, Hal75a, Kel78, Kel20, Nuc20, Pau67, Pie55, Ran55a, Sca11, Van61, Woo61, Ano78m, Pei80, Pog54]. earthly [Her87]. Earthquake [Ano70e]. Earthquakes [Dav79b, Ano92s]. Easing [HS13, Mil50]. East [Ard93b, Bir87, Bro57b, Cav57, Iri79, McL53, Mon53b, Sol80a, Ten75, AS16, Ano84e, Bad66, Bar10a, Ben02, Ber65, Bin67, Bin84, Blu84b, Bou11, Bro16, Brz82, CAM20, Dro16b, Evr78, Fal79, Fel78a, Fel81i, For70, Foz61b, Hal93b, Har14b, Hay79, Hol85a, HGV10, IS15, Jon90b, KL67, Ken81, Khr63, Kla91, Kla97a, Kla97b, Kom90, KJE16, Kum95, Lan14, Mem79, McC88b, ML88, Pea99, Pol67, RM16, Ram04, Sai14, Sak84, Sal79a, Sch58a, SS89, SW59, Sut97, Top58a, Was66, Wil86b, vL85, vMK88, Jon90b, Mye89]. Easter [Fis83, Fis84]. Eastern [Con88, Gol93d, Buk97a, Fis90, KAT^{+}71, Lar89, Lon92, Mak06, Mar68, Nel90, Oli48, Tis90, TM97, Wal50]. Easy [Moo99d, Pra04b, Ano87-34, HS13]. Eat [Ahm00a, Siv75, Ano95n]. Eatherly [Wai60]. Eating [Moo93e, DeW83]. Eaton [Pol79a, Rab60e]. Ebola [SAB^{+}16]. ECA [Por51]. Echo [Ano61j]. Echoes [Isa05b]. Eckert [Smi58e]. Eckhardt [Lyt79]. Eckholm [Han75, Ano78k]. Eco [Ano78n]. Eco-Fiction [Ano78n]. Ecocide [Gof71, Pal72, Wes74a, Dow82b, McC93]. Ecological [Ehr81, Kri85, Smi82a, Gro84b, Wes71]. Ecologist [Hul78]. Ecology [Ano90o, Bur72, Cle70, CH71, HE72, Nel74a, dJ57a, Cas77g, Wil73a, Lew72b]. Economic [Abr74, Ada86a, Ano75c, Ano78o, Ano88e, Bau60, Bra09, Bro79b, Bro67b, Cam89, Car69, Cav53, Col74, Cru67, Dal86, Der66, Der69, Dew55, Don66, DG86, Fis62c, Gaf65, Gol88g, Gre62, Hoo63, Hud60, IW49, Joh65b, Kat53a, Khr57, Kol63, Lew56, Mah59a, Mal54, Mar46a, Mei54, Nel57, Net70, Nor74, Pal64b, Pat62, Pla57b, Pot70, Pre60, Rep85a, Sch69a, Sch72d, Sch47, Soh64, Sol66, Str67, Wil67c, Wou81, Joh81b, Per78, Sea85b, TS22, vD78, Bea86a, Lof51]. Economical [Mei61, WW61]. Economics [All86, Bar79b, Bou61, CH71, Der69, Gra69, Har64a, Hou76, Lan85, Leo65, Mar46b, Mec63, Mil88, Nat63, Pie60, Pri86b, Sch65b, Sil64a, Sle84, Col84a, Gal81a, Lov13, Mar89a, Mat12, Nor11, Ram13b, Sch13b, Sch80d, Sch83, Van90b, Mar54d]. Economist [Eis68, Har47, Ler48, Sap80a]. Economy [Ano66e, Ano69i, Ber93a, Ber75c, Bro56c, Chu04, Coo56, Dea51, Dun66, Fel74b, Fow58, Gol95d, Gro53a, Jon65, Ler49a, Lew56, Nel74a, Orr56b, She74a, Van90b, Wei56a, dJ57b, dJ57a, Ben76b, Cra82b, DeG84, Dum86, FTvHW76a, Gol91a, Gol82c, Kis89, Kla76, Klu22, Rep85b, Sob21, Tan84, Gol93a]. Ecoscience [Eck78]. Ecosystem [PS72, Sch72c, MD24, Mer24]. Ecosystems [Por71]. Ecotheology [Whi72]. Ecuador [Fly05a]. ed [Nor02a]. Eddies [Sut68b]. Eden [Eba72, Woo05]. Edgar [Mul66, Pow89b, The89e]. Edge [Khr99a, AM21, Che20, Mer24]. Edged [CGO05, Wei80b]. Edited [All55a, Ber79, A.54, Abb84a, Ada61, All56l, And65, Ano78f, Ano78n, Ano78l, Ano78m, Ano81c, Ano85d, Ano85b, Ano85c, Ano86k, Ano86i, Ano87p, Ano87o, Ano87l, Ano88e, Ano90o, Ano91s, Ano91r, Ano98h, Bal79c, Bam85, Bar66a, Bas80, Bei84, Ber77b, BD79, Bla84, BGR61, BB63, Boy05, Bre93, Bri86, Bro54, Cav58, Cha74, Cra83a, Cra84b, Cra84c, Cra88, Ded62a, Deu83b, Din69, Div79, Don86b, Don86a, Dow81a, Dre82b, EE85, Eli85, Eva86, Fei77, Fel68b, Fel81a, Fer53, Fis83, For87, For80, Fra61a, Fri64, Gab73, Gil83, H.84b, H.85, Har79b, Her88b, Hob85, Ing82, Ing79a, Jac80b, Joh51, Kat52b, Kat54d, Kat54c, KFM50, Kin86, Kre83a, Kre85a, Kri85, Lac85, LeR50a, Leb94, Lyn69, McK74, Mes64b, Met83, Mil86b, Mol62a, Moo93b, Moo98b, Mor62d]. edited [Mye87a, Mye88, Mye89, Nei82a, Nor64, Par86, Pat79a, Pol68, PF75, Pre96, Rea80, Rei78, RK89a, Roc81, Rot97b, S.80, Sag86, Saw79, Saw81b, Sch79b, Sei60, Sha50, Sie85, Sig85b, Sim85, Sin67a, Soc54, Som84c, Sta85, Ste95, Sti83, Sto59b, Sut54, Swe86, Tho80d, Tho80g, Tic94, Tod80, Tsi72a, Vuc69, Wal85a, Wel79, Whe84b, Wil81a, Win84b, dF79f, dF79h, dF79g, dF80]. Edith [Bal79c, Ano46-80]. editing [Mec20f]. edition [Don89, Gar83a, Mol62a, Tra71, Ano89e]. editions [She85e]. Editor [Ano79e, Day77e, Fel77k, Fel80k, Moo92b, Moo92e, Moo92d, Moo92c, Moo92f, Moo92g, Moo92i, Moo92h, Moo92j, Moo93c, Moo93d, Moo93e, Moo93f, Moo93g, Moo93h, Moo93i, Moo93j, Moo93k, Moo94c, Moo94d, Moo94e, Moo94f, Moo94g, Moo94h, Moo95f, Moo95h, Moo95g, Moo95i, Moo95j, Moo95k, Moo96k, Moo96h, Moo96i, Moo96g, Moo96j, Moo96f, Moo97g, Moo97h, Moo97j, Moo97i, Moo98d, Moo98e, Moo98a, Moo98f, Moo98g, Moo98h, Mye93a, Qua80, The89a, Tsi84b, Boa91, Cra82d, Woo85, AG63, Alb54, Alb57, Alb64, Ale60, All65, All57k, Amd65, Ame66, And58, AL62, Ano53o, Ano66f, Ano66g, Ano66h, Ano66i, Ano67e, Ano75a, Ano80m, Ano85l, Ano16b, AH62, Apt64, Arm53, Arn54a, Arp61, Ash60, Asi57, Asi62b, Aud65, Aug56a, Aus60, Bae60, Bal60b, Bar61b, Bar65, Bar64b, Bat52, BPR52, Bec62, Ben61, Ber66b, BE62, BS67]. Editor [Bet60, Bey46, Ble54, Blo66, Boe52, Boe57b, BR57, Bro66, Bro47e, BGR60, BH65, BJH68, Bur50, BTCSC^{+}66, Cal57, Cam53, Cam60, Cha64, Chi59b, Cla54, Cla59, Coh60, Coh62, CS63, Con60, Coo58, CMMF62, Cra82c, Cra83d, Cra61, Cro16a, Cro16b, Cro16c, Cro16d, Cro17c, Cro17a, Cro17b, Cut58, Daf64, Dav57a, Dav55, Dav65, DFJ^{+}63, Day75j, DeN65, DeV62, Den47, Dix66, DFF^{+}63, Don57, Don64, Doo64, Doo60, Dou60, Doy53, Dre46, Dub57a, Dur69, Dys59, Eds62, Elm66, Elw64, Eme54, Ewi60, Fel62b, Fel66c, Fel78k, Fel78l, Fel78m, Fel66d, Fel57, Fer57c, Fer60, Fin57c, FL62, Fla60, Fle59, Fle58, Fol60, FW64, For64, Fow55, Fow58, Fow59, Fox52, Fra65, Fre57, Fry58, Ger58, GF64, Gil65, Gla63, Gle52, Gom62a, Gom62b, GG63, Gou59, Gou61, Gou63]. Editor [Gra61, Gre47, GC63, Gre63b, GW65, Hah61, Hai54, HFO63, Har61, Har59, Har60b, Har64b, Har65, Har62, Heb52, HH65, Hig63, Hin60, Hla65, Hoc57, HR51, Hol04, Hoo63, Hor60, How63, Hub57, Hub60, Hud60, Ich64, Ing64, Ivy64, Jeh60a, Jeh60b, JvL63, Joh64a, JC54, Jon63a, Jor67, JR62, Kas60, Kat54b, Kaz61, Kea60, Kee56, Kel62, Key57, KH63, Kin67, Kir55, Kis54, Kli59, Koc57, Kor64, Kos60, Kra57, Kue61, Kuh66, La 66, Lap60a, Lau60, Lau64, Led66b, Lee60, Len60, Lev60, Lew61a, LS65, Lew72b, Lie62a, Lin55, LRD^{+}91, Lon52b, Lor57, Lyf61, LM61a, Mac54a, Mal65, Man66a, Man57, Mar54c, Mar54b, Mar58c, MFC62, Mar53c, Ma’52, Mas57, McA65, McC54b, McC54c, McC57, McI69, McM58, Mec16a, Mei51a, Mei61, Mer54a, Mer54b, Mil62]. Editor [Mil63a, Mil63b, Mit58, Mon53a, Moo97f, Moo97e, Moo00e, Mor57a, Mor62a, MD66, Mor66a, Mor61b, Mos65, Mot64, Mun65, Nea61, Nor56, Ols62c, Olt52, Opp62b, Ora65, Ore62a, Ore62b, Orm57, Pac59, Paj52, Par66c, Pas51, Pas66, PR62, PP63, PWW63, Pfe52, Pip55, Pis64, Pit52a, Pol57a, Pol62a, Pol62b, Pon65, Pot49, Pri58b, Rab48d, Rab56d, Rab56f, Rab59b, Rab59c, Rab60e, Rab61j, Rab66b, Rab66a, Rab67d, Rab67c, Rab67e, Rat62, RMK65, Rin61, Rin66, Rob49, Ros62, Rum54, Rus50, RGR63, RB65, Sac57, Sch66, Sch54b, Sch57a, Sch04b, Sch63b, Sch63d, Sco66, Sea56, Sel59, Sen65b, Sha65, Sha66, She64b, She65, Shi66, SdSP62, Sie52, Sil65, SIF66, Sin63b, Smy64, Soc55, SD56, Sol66, Spe54, Spi66c, Spi67, Spi63, Sta60a, Sta61, SJ64]. Editor [Sti62, Sti52, Sti57, Sto62, Sto60, SH61, SG59, SG62a, SG62b, SG65, Szi60b, Szi61a, Tay52, Tay56, Tay62, Tay69, TK66, Ten69, TW66, Til64, Tod55, Tsi78a, Tsi83h, Tsi83g, Tuc62, Tuc64, Tuf54, UC62, Ure56, Ure61, Vis59, Wal55, War59, Was63, WW61, Wei46, Wes58, Whi53, Wie47a, Wig58, Wil62, Wis55, Wit52a, Wit52b, Woo62, Wor58, You64b, You69, Zar64, Zei60, Zel70, Zif53, Zlo47, deF51]. Editor-in-Chief [Ano75a, Ano85l]. Editorial [Ack90a, Ano50s, Ano51k, Ano63l, Ano63m, Ano63n, Ano64j, Ano84e, Ano91-62, Ano16a, Ano17c, Bac81, Bet81a, BL83b, Edi54, Coc51, Coo75a, Day74b, Day74c, Day74d, Day74e, Day75d, Day75i, Day75f, Day75g, Day75h, Day75e, Day76b, Day76c, Day76d, Dje75, Dub50a, Fel67b, Fel71, Fel75c, Fel76c, Fel76d, Fel76b, Fel76e, Fel76f, Fel76g, Fel76h, Fel77c, Fel77b, Fel77d, Fel77e, Fel77f, Fel77g, Fel77h, Fel77i, Fel77a, Fel77j, Fel78c, Fel78i, Fel78e, Fel78f, Fel78g, Fel78h, Fel78d, Fel78j, Fel79i, Fel79d, Fel79e, Fel79f, Fel79j, Fel79g, Fel79c, Fel79h, Fel79l, Fel79k, Fel79m, Fel80c, Fel80d, Fel80e, Fel80b, Fel80g, Fel80f, Fel80h, Fel80j, Fel80i, Fel81f, Fel81e, Fel81h, Fel81i, Fel81d, Fel81j, Fel81g, Fel81k, Fel81l, Fel82d, Fel82c, Fel82b, Fel83b, Fel84c, Fel84d, Fel84e, Hal83, Has67, Lap53, Lew67c, LM76, Lon75, Myr74, Rab46c, Rab46d, Rab46e, Rab46f]. Editorial [Rab47a, Rab47h, Rab47b, Rab47c, Rab47i, Rab47d, Rab47e, Rab47j, Rab47g, Rab47k, Rab47f, Rab47l, Rab47m, Rab48f, Rab48g, Rab48e, Rab49c, Rab50b, Rab50d, Rab50c, Rab51b, Rab51c, Rab51f, Rab51g, Rab51e, Rab51d, Rab51h, Rab51i, Rab51j, Rab52b, Rab52c, Rab53c, Rab54c, Rab54d, Rab54e, Rab54a, Rab54b, Rab54f, Rab55a, Rab55b, Rab56b, Rab56c, Rab57d, Rab62c, Rab70d, Rat71b, Sap75a, Shi52, SG75, Wal74, Zim48, vH75a, ATL85, Bro62b, Dre87, Fel74a, Fel93, Gla60, Ing55a, Ing74, Mur51, Myr80, Osb54, Pei53, Pri74, SL87, Sup61, Tuf54, Wil90, Ack88a, Cra82b, The88, The90]. editorials [GF79, Rab56d]. Editors [Lan56b, Ano96n, Coh95a, Leb91b, Wes91, Tsi83c, Ack87g, Ack87f, Ack87c, Ack87e, Ack87d, Ack87b, Ack87h, Ack87i, Ack88d, Ack88b, Ack88c, Ack88k, Ack88f, Ack88e, Ack88g, Ack88i, Ack88h, Ack88j, Ack88l, Ack88m, Ack89a, Ack89g, Ack89b, Ack89d, Ack89e, Ack89c, Ack89f, Ack89h, Ack90b, Ack90c, Ack90d, Ack90e, Ack90g, Ack90f, Ack90i, Ack90h, Ack90j, Ack90k, Ack91b, Ack91c, Ack91d, Ack91a, Ack91e, Ano48e, Ano85k, Ano85r, Ano85l, Ano85n, Ano85m, Ano85o, Ano85p, Ano85q, Ano85s, Ano85j, Ano86y, Ano86z, Ano86-27, Ano86x, Ano86-30, Ano86-29, Ano86-28, Ano86-31, Ano86-32, Ano87z, Ano87-27, Bro85b, Bro85c, Bro85d, Bro85e, Bro85g, Bro85h, Bro85f, Bro86b, Bro86c, Mar89f, Mar48e, Moo91b, Moo91e, Moo91c, Moo91f, Moo91d, Moo92k, Szi47c, Szi61a, Wei52a, Han98a]. Edmund [Hol83a, Lit54]. eds [Ano10a]. Edson [Gre68b]. Eduard [Han75, Sho92b]. Educating [Ano88e, Aus81, Pri68, Ano88e]. Education [AW84, All59l, Ano48f, Ano66q, Ano88e, Ast67, Bas68, Bou82, Bur66, Con64, Dav58, DuB66, For70, Foz61h, Foz62d, Fri56, Gen86, Gil72, Gla58, GS68, How56, Jam58, Kal82c, Lap73b, Mey56, Old66, Orl84, Pie66, RS58, Ric67, Rot58b, Str71, Sup67, Thi59, Bal82, BRB16, CN88, FWCL20, Hes84, La 88a, MDB81, NR85, Rap11, Rin84d, Sim84a, You85b, Bea47b, Kat57b, Lof47, Ros67, Ker68, Pla69, Wil53b]. Educational [Eme47, Noy46b, Smi71, Wol58b]. Educator [Fel92]. Educators [Sim47].

Edward [And65, Ano85c, Ber78a, Bie68, Bur90c, Dow83b, Gin54, Hir88, Hyn69, Lin85, Mil86b, Mor62b, Ore58a, Pal02, Pee78, Pei80, Per87b, Rei74, Roc85, Rot97b, Rus51a, Sch80c, Tho80g, Woo73b, Wri63a, vH83b, And72, Ano55i, Ano89r, Ano90-31, Ano95q, Ano02g, BO76, Bul83, Con55, Gaf86, Goo04a, Goo04b, Har10, Jac74c, McM03, Pub55, vH83a, Ano05b, Her90b]. Edwards [Pat81a, Wil81a]. Edwin [Ano78n, Cow65, Sol68, dF79h]. Effect [Ano46-71, Ano48u, Bru57, Lew68e, Vel84, Ano94m, Kun23, Win81a]. Effective [Ano54x, Ber55, Cal80, Elm66, Mil86b, Rid50b]. Effectiveness [Cal73, Kin80b, Kno66]. Effects [Ano46q, Ano49f, Ano56h, Ano56n, Ano61k, Ano78f, Ant70, Arn54b, Bon51, Bro63a, Bru48, Bru58, Cog82, Coh76a, Cro58, Has57, Hir50, Kap66, Lal65a, Lap48, Lee73, Lib58, Mol62a, Mul47, Neu58a, Neu58b, Pau62, Pau63, Pau70, Rum71b, Rus56b, RGR63, TG70, Upt63, Wal51, WH72, Ano15a, Bar77b, Due21, Ehr84, Gre12, GCV16, Hil12, Mac81b, Nat77, NBM85, WvH86, vHC86, McC96, Pre96, Lap56a]. Efficiency [Kre77, RW76]. Efficient [Pap88, Rub74]. Effort [Ing57a, Nat74, Shu78, Ano86f, Bar77b, Wil92c]. Efforts [Pau52b, Rab60a, Sch62c, Bul08b, KM90, Kor19, Sal16, Vil75]. eggshells [NPS90]. Egon [Smi54a]. Egregious [Rot95j]. Egypt [All60h, Dil57, Fah90, Kum95]. Egyptian [Sai14]. Ehrenreich [Kel79]. Ehrlich [Eck78, Smi85d, Ano11e, But73, Har81, MW72, Sch86]. Ehteshami [Jon90b]. Eibesfeldt [Tho80g]. Eibl [Tho80g]. Eibl-Eibesfeldt [Tho80g]. Eight [BCH^{+}48, Lap53, Rot95j, Dow82c]. Eighteen [Got83]. eighth [Isa83b]. Eileen [Sto18b]. Einführung [Thi46]. Einhorn [Kin86]. Einige [Ano65e]. Einstein [For80, Rot81b, Sha54b, Ano48x, Ano48-33, Ano92-32, Bro79c, Dow80g, Ein63, Fel79n, For80, Har79c, Isa07, Kol07, Mar79c, Opp79a, Opp79b, RF55, RF79, Rot79b, Sea79, VFIS48, VFIS63, For80, Kap61]. Einstein. [For80]. Eisei [Cog82]. Eisenhower [Gol78a, Hor77, LaF86, Ano54-29, Ano55q, Ano58d, Bon53b, Eis57, Eis61, LPT^{+}54, Rab55a, Sim54d, Str05c]. Ekirch [Woo65]. Elastic [Com75c]. Elasticity [Moo97h]. ElBaradei [HP09, Bul09f]. Eleazar [Pla59a]. Election [Cla82, Gom80, Isa82b, Kal82d, Mar82a, Pai82c, RF65, RR68, Bar79c, Isa84a, Mec16e, Mec24, Sto24, Swe83, Khr96b]. Elections [Isa02d, Isa03e, Isa04d, Ano48u, Gau24, Law90, MD24, YWS20]. Electric [Ano51l, Ben71a, Cis52, Jon71, Lon78, Mil71, Net69, Net70, Spe18, Spi72a, Wou71, Ano49e, Wou81]. Electrical [YY74]. Electricity [Lit54, MA83, Ano11h, CR17, DL21, Pan92, Ram13b, Ano83l]. Electromagnetic [Har91a, Ste83a]. electronic [Pla78]. Elegant [Bos78, Per88a]. Element [Fly00a, Ano97i]. elemental [Hec12, Hec20]. elementaren [Thi46]. elementary [Thi46, Ano57j]. Elements [Hol75]. Elephant [Cav13, Isa04e, O’L73, Abb91, Hud78]. Eleventh [Kel98a, Tuc97, Ano85r, Ano63o]. ELF [NA89a]. ELFs [Jau90a, Jau90b, Sha90]. Elie [For80]. Eliminate [Oel12, Ber81, Cho81, MDB81, Shu81, Ste81a]. Eliminating [Maw76, Tal18]. Elimination [JC50, HNR12]. Elis [Ano57c]. Elite [Nad79, Sal75b, Shi51, Zim75, Bei82, Cra82c, Mow95b, Pop93, Fal61b]. Elizabeth [Ing70, Ano14b, Dro24e]. Elkana [dF79f]. Ellen [Eva86, Whe84b]. Ellis [Sim88]. Elmer [Wei55b]. else [Dro23e, Fie21b, Pat78b, Wea82]. Elting [Ada67]. elusive [Pai82d]. Emanuel [Sto17b]. emasculates [Smi98]. Embalming [Ano95k]. Embargo [Ano99n, Ger82, Con86]. Embarrassed [Ano95r]. Embarrassment [Kum99, Day92, Ust93b]. Embrace [Kre90]. Embracing [Kra06]. Embryology [Bro56a, Fra72]. Emelyanov [Rab67a]. emerge [Hec22]. Emergence [Gou70]. Emergency [Bey80b, Hir55, TB23, Gra85, GT14, SR81b, Ein49, Fer68, RG49]. Emerges [Pag68, Ack89b]. Emerging [All15a, Cra88, Mec23c, Pol09a, Abr86a, Abr86b, Ame86, Ano11f, Ber86a, BG14, DH86, Esp14, Fis86b, Hof86, Hun14, MA17, McC86b, McC86a, Pat86c, Tho86a, Tom14, Vog13, Wei86b, Wei86a, vHC86]. Emigration [Git74, Mar72b, Rab75, Gor85]. Emigré [Git74]. Emile [Dow83f]. Emilio [Wil71]. Emily [Dro23m]. Eminent [EBU^{+}52]. Emirates [Al 11]. Emission [Spi71c]. emissions [Boo22, Cav13, Coh76d, JP13, Kam20]. Emotional [Oli58, WW58]. Emotionalism [Bru55b]. Emotions [Eis66]. emperor [Fel79m, Que81]. Emphasis [Kep76]. emphasize [TS22]. Empire [Ano90o, Wil87]. Empleado [Ano91u]. Employee [Ano53y, Fel83d]. Employees [Bur50]. employing [Ken71]. Employment [Ano55j, Ano55w, BS47, Bro55c, Gou47a, Gre58, Lil48]. Empty [Cav05, Sch05g, Wou81, Moo95q, RC87, Ano86i]. emptying [Kas77]. ’ems [Coh87]. enabled [Lin19]. enablers [Gif13]. enacts [DW85]. encounter [Bro79c]. Encounters [Ano02g, Hei70, Hei71]. Encouragement [Opp51c]. Encyclical [Pop52]. Encyclopedia [Mye88]. END [Gom81a, Ano72e, Ano99-27, Ark99a, Bro47d, CB95, Dow80e, Fel83c, Fel84c, For91, Gol96a, Hum72, Kal88, Kim01, Lew68d, Lew71c, Mac91, Mar65i, McC57, Moo97a, NA90s, NA92l, NA94d, NA96f, NA96d, NA97e, NA97a, NA98g, NA98d, NA99c, NA99d, O’L75, Ols62b, Rab63h, Rob91b, Sch57b, Sha98b, Sin67b, Smi85c, Uhl95a, Wil75c, Wil84, Wol91a, Ano92t, Ano99-29, But97a, Cas19, ES89, Fai97, Fal87, Fra89, Hun90, Moo92d, NA88s, NA89n, NA90o, NA91n, NA92c, NA93n, NA93r, NA94h, NA95f, Smi92a, Swe79a, Wei94c, Ber80, Fel81h, Fre81, Kal81, Tho81c, You81]. Endangered [Ada87, Bal79c, Buc77, Fra72, Mea72, Wil77a]. Endangering [Bal79c]. ENDC [Jac67]. Endeavor [Fei07, Ric74]. ended [McK23b]. Ending [Bak06, BM99, HM15, Mur60, Uhl97b, WMT93, Ano86h, Sha95b]. Endless [DuB57b, Sha87b, Tay86, Tho80d, Woo99a, KN14d, Ano98g]. Endorsement [Ano52s]. Endorses [Tru46, Gol90c]. Endowment [FCC^{+}46]. Ends [Cav55, Wig61, Ano14b, Cum97b, Nel92a, WM78]. Endurance [dF79g]. Enduring [Str07b, VA11, DM24a]. Endy [Sie07b]. Enemies [Ano97m, Sti67, Ano91-43, Ano92-28, Moo95b, Sag92]. Enemy [Ano51h, BF05a, Day97, Lue96, Ano93u, Ano96j, Ark83a, Fel78f, Pet92]. energies [Smi79, SS80]. energize [Gya14, Har14a, Sha14]. Energy [A.54, Ada52, Aeb48, Alf72, Alf74, All55l, Alv00, And65, And71, Ano45e, Ano45f, Ano45o, Ano45p, Ano46j, Ano46l, Ano46m, Ano46-104, Ano46-101, Ano46-102, Ano46-103, Ano46p, Ano46w, Ano46y, Ano46z, Ano46-45, Ano46-32, Ano46-106, Ano46-50, Ano46-60, Ano46-66, Ano46-93, Ano46-95, Ano46-92, Ano46-90, Ano46-105, Ano47c, Ano47b, Ano47d, Ano47e, Ano47f, Ano47j, Ano47r, Ano47w, Ano47y, Ano47-27, Ano47z, Ano48g, Ano48i, Ano48j, Ano48k, Ano48l, Ano48m, Ano48n, Ano48h, Ano48w, Ano48-28, Ano48-29, Ano48-31, Ano48-34, Ano48-39, Ano48-41, Ano49-30, Ano49-39, Ano50l, Ano50v, Ano50-35, Ano51l, Ano51r, Ano51-28, Ano52b, Ano52-27, Ano53a, Ano53b, Ano53c, Ano53-43, Ano54o, Ano54q, Ano54t, Ano54u, Ano55l, Ano57d, Ano60f, Ano71c, Ano78d, Ano78o, Ant70, AM74, Arn47]. Energy [Bac48, Bar74, Bar73b, Bas80, Bec58, BAB^{+}46, BH62, BF73, Ber75c, Bet77b, Bey46, Bha55, Bla46, BBD^{+}46, BO76, Bro57a, Bro80a, Cou48, Car49b, Car49c, Car49d, Car49e, Car50a, Car50b, Car50c, Car50d, Car50e, Car50f, Cav50, Coc50, Coc55, Coc59, Coc51, CH54a, Col59, Com57a, Coo81, Cre74, DB52a, DB52b, DB52c, DS47b, Dav72, De 74, Dea51, DR71, Don89, DPL75, Dow80f, Dro16a, Eme71a, Eme71b, EBC83, Fel74a, Fel77c, Fel83e, Fet00, Fis46, Fol74, For09, Fox46, Fox48, Foz61a, Foz62b, Fra61a, Fre71, Fre73, Gee71, Gla75b, GB73, Gol62a, Gol62b, Gor81, GS09, Gre62, Gre67b, GW02, Gro47c, Gué70, H.84a, Haf51, HW73, Han46, Har50b, Hei47, Her78a, Hin53, Hir73, Hod55, Hoe88, Hol75, Hou76, HM72, Ing71, Jac74b, Jol46]. Energy [Jon59, Jos61, Kal54, Kal63b, Kat57a, Kel79, KB80, Kih47a, Kih47b, Kih47c, Kih47d, Kih47f, Kih47e, Kih48a, Kih48b, Kih48c, Kih48d, Kih48e, Kih48f, Kih48g, Kih48h, Kih48i, Kih48j, Kih49, Kli81, Kow48, Kra73, Kra59, KDFS81, Kuy54, Lak82, Lap57, Lau47, LeR49b, LeR50b, Lei47, Ler49a, Lev46, Lev47, Lev48, Lew50, Lib59b, Lie46, Lil46, Lil47a, Lil47b, Lor81, Lov74b, Lov74c, Lun59, Mar48b, MT55, Mar48c, Mar70, Mat97a, McM48a, McM67, MM71, Mer72, Mer79, MR09, Mil48b, Mil52, Min48, Mor82a, Mos81b, Mur54c, Nef78, Nel84, Net71, Nie63, Noy46a, Ols74a, Oom46a, Oom46b, Opp46, Opp48a, Opp63c, Osb49b, Par71, Pat78a, Pat81a, Pat82a, Pau70, Per55, Pic51, PD12, Pol68, Por73, Pri56, Rab47n, Rab47a, Rab47d, Rab47k, Rab48f, Rab48e, Rab49a]. Energy [Rab49c, Rab52a, RS52, Ran53, Ran55a, Ran59, Red78a, Red78b, Rex71, Rho46, Ros08, RW76, Rot79a, Rub74, Rub73, RL78b, Sch78a, Sch47, Sea65, See79, Shi46a, Shi47c, Sim54d, Sim54e, Ski53, Smi54c, Smi74b, Smi75a, Smy50b, Soc54, Sør76, Spe49, Spe89c, Spi65, Spi71a, Spi75, Spo50, Spo51, Bul08r, Bul09d, Sta73, Ste52, Sti83, Sto72, Sto77, Str58, Szi50a, Tel51a, Tel48b, Tho53, Tra73, Tur51, Tur52, Ung78a, Ure46b, Ure47, Ure48a, Usm79, Van46, Van52, War48, Wea79, WH00, Wei70b, WH72, Wei56b, Whi71, Wig51, Woh68, Wou81, Wri46, LLR^{+}47, You76a, Zim77, Zlo47, dR49, vHW76b, vHW76a, vN57, AH79, AGB22, Ano47k, Ano12b, Ano12e, Ant76, Ant78, Avi83, Bac77a, Bar77a, BW80, BPTB76a, Bis14, Cap77, Cas77a, Cas77g, Coo78, DR14, Day75a]. energy [DJ13, DL21, Dro16b, Dro21d, Eng78, Fed78, Fel77b, Fel14, Fer10, Fot89, Fre76a, Fre76b, Fri21, Gho15a, Gol78b, Gol79, GG11a, Got77, Gow79, Hai81, Har79a, Hec19, Hec16, Hen78a, HS97, Hib12, Hol77, Ikl76, IES79, JP13, JFS20, Kam14, Kas15, KJE16, Kre81, Lan90k, Lif76, Lil62, Mac86b, Mak76, Mak24, May17, Mez12, Mor14, Mor77, Mye91, NF13, Not15, Pal16, PW76a, Rab50e, Rab16, Rem15, Rom91, Ros78a, Rot78a, Sak78, Sal80b, Sch13a, Sch13b, Sel77, Sie82, Sin83, Smi77a, SS76, Squ21a, Sto17b, Tak11, Tay81, Tel47b, Tho80a, Twi23, War14, WW60, Wei77a, Wei80a, Wil79, WP80, Wol76, Zha15b, Ano47t, Ano47u, Ano56j, Ano60f, Ano91-41, Ano91-28, Ano91-32, Ano94p, Ano97m, Bas80, Bro23, Dro23a, Hoe87, Lan90c, Mak94a]. Energy [Mor85b, Nel93b, NA91g, Pat81a, Pos79, RDS^{+}52, Ram88a, Wee98, All55a, Bru50, Cha58b, Dow80d, Dow83a, Fel83a, Ham61b, KFM50, Lap57, Pos79, Sha50, Bas80, Ber78a, Cha83b, Fel84b, Fre76b, Kat54f, Nel74a, Pat81a, Pei80, Sch78a]. Energy-Limited [Ber75c]. enforcement [Nay12]. Enforcing [Wei61]. Engage [Uhl05]. Engagement [Zah06, Isa88f, Len11, Pep15, SD17]. Engaging [CYP14, Sto18b]. Engelhardt [Saw79, Rot97b]. Engender [Bae60]. Engineer [Alb93c, Ano98g, Dru71, Dru72, Fel92, Lev72b, Bar81b]. Engineering [Ano46-56, Ano54y, Cha69c, Dan72, Dix66, Dje72a, Edm75, Fre72b, GT14, Hoa72, Jai73, Kat57b, Kil62, Nag71, PB68, Ram72, Rei50, Sta00, Tin67, Wei66b, Whi51a, CYP14, Cav82, CO14, DK80, Pal84, GC83, All54d, Sin67a]. Engineers [Ken71, DS47b, YG77a, Kat57b, Sea56, Yar80a]. England [Ack97, Blo95, Boo46, Gla75d, Sim46b]. Englebert [Ano84f]. English [McK74, Jam58, Mar49a, Mas48b, Tam71a]. Enhance [SD09, Kok88a]. Enhanced [Mie77, Agn77b, FG14]. enigma [Ber04c]. Eniwetok [Ano51m, Str56]. enjoying [Isa97c]. Enlightenment [Man72]. Enmities [Nel93a]. Enola [Rot97c, Ano95s, Har96]. enormity [Ste77b]. Enough [Bla85a, Dav77b, Dia83, Dow83a, Gul75, HF80, Hot75, MF68, Que67, Rei02, Ric89, Ano95n, Bir89, Car03, EG89, Hir89a, Moo95p, Ray90, Rot77, Sam18, Sch11, WC17]. Enquiry [Ano45c]. Enrage [Uhl05]. Enriched [Ano92r, Rot93a, Moo15, NA94e, SD10b]. Enrichment [BF05b, Gla76a, Kwo10, AS16, AH91h, Cas77b, Dic77, Shi88]. Enrico [ASA55, Ano68g, Seg70a, Seg70b, Wil71]. Enroll [Men92a]. ensure [Kib12, Nay12, Squ12]. Entangled [Edw12]. entanglement [Mak24]. Entangling [Pow90a]. Entebbe [Day76h, Med76]. Entente [AB05]. Enter [Ano85k, Hec22, Lop78, Whi89]. Entering [Smi90a]. Enterprise [Ano53a, Ano53b, Ano53c, Ano53d, Ano53-43, Ano63v, Rin75, Sag73, Ski53, Tod64a, CZ95, Har14c, Wei76a]. enterprises [Kow77]. enters [Ano94p, Mak94a]. entirely [DJ13]. entrenched [NH08b]. entropists [MW83]. entropy [Wei82a]. entwined [Sch19]. Environment [Alf72, Ano72h, Ano78n, Aur72d, Deu91, EH72b, Eli71, Fed72, Fel79e, Gle91, HM72, ILA70, Lee73, LR71, Mil02, Pec70, Pit61, Wei70a, Wei70b, Wes74a, Ano78p, Ano15g, Cla82, Eps77b, Fou84, Mye87b, RS88c, Rob90b, Sut84, Ano78n, PF75, dF79f, Ack97, Eck78, Rob69, Sin76]. Environmental [Abr74, Ano70e, Ano71g, Ano78g, Ano78k, Ano78n, Ano82f, Ano90o, Ant69, Ant70, Bar76c, Bat73, Ber72c, Ber73c, Bye91, Coh76d, Col74, D’A70, Gle94, Har72b, Hol80a, Jac72b, Lew70e, McC93, Mic74b, Mye87a, Mye93b, Nel74a, PE24, Pre96, RR81a, Sch90, Sin81, Sta00, Tra73, Wou81, Yar80a, YY74, vH77a, Ano11c, Ano11e, Ano14f, Ber77b, Dro19, Koe22, Kri85, PF75, Voh71a, Ano91q]. envoy [Pal77a]. Envy [SW07]. EPA [Ano73h, Ano95j, Por74]. Epic [Ano72e]. Epidemic [Dow81h, Smi05, FWCL20, Hia80, Dow82h]. Epidemics [Kap60]. Epigram [Tra51]. Epilogue [Her82a, Ney73]. episode [Mag88a]. Epoch [WSG^{+}15]. Eppur [Rab61d]. Epstein [Ing82, dF79c, Ano86j]. Equal [dF80, Ano11h, GK21, Sub93]. equal-opportunity [Sub93]. Equalizer [Res76]. Equation [Har69b, Jac80b, Oue07]. Equilibrium [Cha74, Ben76b]. Equipment [Ste69b]. Era [Ark99a, Bro67a, Gre67c, Lew71c, Mur60, RR68, Squ75, Tel71, Tél68, Ano91-62, Ano91-65, Cro17d, DW90b, Fel83c, Fie18, Gle24, Gra82a, Pep15, Leb91b, Ano82k, Moo96c, RK89a, Swe88]. Eradicate [Sub87]. erase [Isa95e]. ERDA [Smi75a]. Erdith [All55b]. Eric [And68, Ano87s, Rab48c, Say84, Bul07d, Tho80g, dF79f, Ano14d, Dro23c]. Erich [Las62]. Erik [Ano78k, Dow80h, Han75]. Erin [Bre74]. erinnert [For80]. Ernest [HI72, Nei82a, Pom66, Rei72]. Ernst [Hur78a]. erodes [Ode88]. Eros [Ano93n]. Erosion [Bal85]. Errant [McA71, Arn22]. Errata [Ano46-39, Ano46-40, Ano64k, Ano74e, Ano81e, Ano82g]. Erratum [Ano48y, Ano49o, Ano53p, Ano53q, Ano54k, Ano55j, Ano66j, Ano80i, Ano80j, Ano82i, Ano82j, Ano82l, Ano82k, Ano82h, Ano83j, Ano83k, Ano83l, Ano83m, Ano84h, Ano84g, Ano86s, Ano86t, Ano86u, Ano87x, Ano11a, Ano11b, Ano16c, Arn55, Gom62a]. Error [Wei75, Ano87-38, Ano90-33, Dug91]. Errors [Sch02b, Sko58, Bal84a, Par88a, Pol88]. Ersatz [CO74]. ERTS [Ber73a]. erupts [Kom90]. Ervin [Mye88]. Erwin [Bur65, Day97, Lue96]. Escalate [STC74]. Escalation [Coo71, Har71, How65a, Ist67, KL67, Ano06q]. Escande [Tod81]. Escape [Ols63, Rab62d]. escaped [O’N89a]. Escaping [Got82]. escrow [Tur98]. Especially [Smi16]. Espionage [Ano50r, Dav51, Pow86, Row88, Eav17b, Han87b, Osb82, Seg13]. ESPying [Ano92v]. Essay [Lap61a, Mar61d, Lan90l, ZGD83, Ano78o, Ano82r, Ber81, Cho81, Gra82b, MDB81, Mye82a, Shu81, Ste81a]. Essays [Ano94h, Ber77b, Pee78, Sim65, dF79g, BM99]. Essence [Str05b, PMRP50]. Essential [Mei51a, Mei51c, Got20]. Essentially [Mor51]. Está [Fly02a]. Establish [Ano48d, Ano53w, Wei53b]. Established [Kol75]. Establishes [Ano46-76]. Establishing [Spe58]. Establishment [Dru71, Dru72, Lev72b, Lov73, Mor62d, San66b, Sco71, Fre76a, Hag77, Sco08b, Coc51, Pic51, Ski48, Abr73, Van87]. Establishments [Aur72c, Ber66a]. Estate [Wei64b, Kal83a]. Estep [dK60, Lap57]. Estes [Ano46-90]. Esther [Lad81, Moo95d]. Estimate [Cra83b, Tam74, Arn47]. Estimated [NA88d, NA89e, NA90b, NA91d, NA92d, NA93d, NA93c, Nor94a, NA94a, NA95b, NA96b, NA97a, Sch47]. Estimates [RSW55, Fla80a, Hol80b]. Estimating [vHC86]. Estonia [LJ17]. Esvelt [Eav18]. Etcetera [Ano94n]. eternal [Ano95j]. Ethel [Mon53a]. Ethic [Can74, Jen74, Sta49, Ber77b, Sin76]. Ethical [Hil52, Lie62a, Lon51, Ma’52, Wil68a, vL48b]. Ethics [AH79, Ano82m, Ano85b, Bac59, Cal82, Cra68, Daf64, EH72b, Gus85, Hay64, Hei64, Lew72a, War87, All88, Bro87b, Har79a, Jon84, Nat87, Saw81a]. Ethiopia [Sto78a]. Ethiopian [Fer69a, Teb85]. Ethnicity [Cle82]. ethos [Lif85]. Etzioni [Cha62, Cow65, O’B75]. EU [Sal16]. Eugene [Ano78o, Bro57a, Fla68, Kal68, KFM50, Lyn78, Phe63, Rei78, Ada97, Ano73c, Ano73d, Ano81m, Elm74, Par73]. Eugenics [Orn67]. Euratom [Gor67b, Gué67, HA85, Hir59, Rab65h]. Eurobomb [FC96, Hib96b]. Eurocentric [Sak88]. Eurodeterrence [FC96]. Eurodeterrent [Mei20]. Euromissiles [Osg83, Ano83n, Dea83, McA84]. Europas [RK87a]. Europe [Ano57d, Ano87l, Ber84a, Bir87, Con88, Dew55, Fis83, Gol93d, RK87a, Sco74, Sha95b, Som91, Swe86, Tay87, Ass80, Ack88i, Aga79, Ano60f, Ano81j, Ark86d, Aro50, Aro51, Asp87, Bar83a, Bar80i, Bar82b, Bir85, Bor82, Bor58b, Bri80a, Bur48, Bur62a, Cal84b, Car83c, Cot84, Cra84b, Cro64, Daa87, Eva87a, Fel82c, Fis84, Fis90, Fly06b, Gle83, Gol67b, Gom75a, Gro83, GBG^{+}81, Gué69, Gué70, HFMH89, Has88, Hol83c, How62, Kal85a, Kal88, Ken85, Kin83, KSS17, Lal64g, Lar89, Lei83b, Lod84, Lon92, LJ17, Lun82, MV04, Mak06, MB98, Mar67b, McM82, Mea85, Mow83, Mul65, Mül90a, Mül93, Mye82a, Myr80, NA87e, NA87i, NA90h, NK04e, NK11, Pat86a, Pol69, Rab54g, Rab56e, RK90, RK91, Sap84, Sch04a, Sen87, Sha93b, Sig87a, Sim84e, Swe77]. Europe [Tee65, Tis90, TFM81, TC83, Tsi83e, Usb52, WB06, Wei54a, Wes85, Yor76a, You65b, Zuc89, Bre87, Kre83a, Smi54a]. European [Ber84a, Dew55, Mul66, PP83, Wil81a, Ack88b, Alb83, Ano76d, BH80, Bor52b, Bre72, Chu50, CY81, Dea88, Deu67, Fel82i, Fer52, FC88, Gol78c, HW96, Jor11, Kub88, Lal67h, Lam85, Led66b, LP08, Lin79, McC86a, Min52, ML88, Mot82, OW65, Pai83c, Sau10, Sig85b, Soh64, Swe86, Yaz88, You64a, vW80]. Europeanization [Bir85, CA84]. Europeans [Bar81e, Fit11, NA90n]. Evacuation [Lin55, Pet54, Ste57b, vHC76, Abr86b, Cav79]. Evaluating [Gar08, McC09, Neu58b, Pol09b, Dah78]. Evaluators [Ano46k]. Evan [Alc63, Shi54c]. Evangelicals [Sto19b]. Evans [Nei82a]. Evasion [Ing60b, Slo68]. Eve [Maz69]. Even [Bec79, Bra47, Cra96, Tho84b, Wei94a, AH91c, Gla85, Sez20, Tro20, WA93, Zag95c]. Event [DL78, Wat82]. Events [Ano50h, Cof62a, Cof62b, Cof62c, Cof62d, Fel62a, Kat61a, Mid73, Rab62d, Rub62, Uhl03b, Ano94x, Ano94y, Ano94z, Ano94-27, Ano95y, Ano95z, Ano95-27, Ano95-28]. eventual [Arn47]. Ever [Ano92s, Bla69a, HW08, Fig23, Hud78, McK22, Ste89a, WW80]. Ever-Ready [HW08]. Ever-Widening [Bla69a]. Everett [Sto59b]. Everlasting [Tho80d]. Every [Ano98k, NA92e, Ano96o, Coo90, Pos21, Sch89b, Sha84a, Ste83b, Vos88, Mos91]. Everyone [Ano98m, Hit03, Rab48h, You92, JH16, Rot18]. Everything [Ano90-32, Ark02b, McK25b]. Everywhere [Lan01, Sto60]. Evgeny [Ano73b]. Evidence [Haw73, Zim89a, Kal82i, Mar83e, Mec20d, Mes85b, Mor88, Pab88, Woh78]. evil [Yuk81]. Evils [Bor66]. Evolution [Cam76, Dan72, Dje72a, Dob64, Fre72b, Fuc70, Hoa69, Hoa72, Led66a, Lew70b, Mic69, Mul47, Mul67, Orn67, Rab72b, Ram72, Roc83a, Hoa77, McL11, Was82, Cro67, Fel84b, Pla64a]. Evolutionists [Hay55]. evolve [CAM20]. Evolving [Mur72, Mec17c, Zys17]. Ewa [Ran85]. Ewald [Rob69]. Ewing [Fel83d, DM24a]. Ex [Ano99a, Sal71]. Ex-Astronaut [Sal71]. Ex-Stasi [Ano99a]. Exact [Haw58b]. Exaggerated [Coh76b, Alb98b, Gol96a, Mar93b]. exaggeration [Loc89a]. Examination [Coh48]. Examining [Fra91]. Example [Sac64, Van90b]. excavation [O’N89a]. Excellence [Lal66f]. excepted [Gol93b]. Excerpt [Ano86v, Ano86w, Man23, Hol87b]. Excerpts [Ano52h, Ano58d, Ano61k, Ano79f, Ano79g, Ano80k, Ano85g, For51a, For51b, Oli66, S^{+}98, You68b, Ano85t, Nat85c]. Excess [BM02]. Excesses [Abr76]. Excessive [Ing60b]. Exchange [Ano46r, Ano62-28, Ano64j, Axe60, Bro64, BJH68, Cut58, Gil55, GJ67, KK63, Mit62, NW70, Pow59, Rab52h, Rab53d, Sea63b, Sem53, SSC68, Sim54d, Vie62, Ano80m, Bil85, Cha85a, CO85, DG86, Wes79b, BMPH18]. Exchanges [CS75, Gra83, WW80]. Exchanging [Dub62a, Kim22]. Excitement [Rus52a]. excludable [Per87a]. Excluding [Per87a]. Exclusion [Kal87]. Exclusive [Ram88b]. excuse [vH80]. Excuses [Aue05a, Rot96d, Ark96b, Ark96f]. Execution [Mon53a]. Executioner [Gla75a]. Executive [Ano53-42, Ano46-65, Ano55w]. Exemption [Mur51]. Exercise [Gus95, Whi78a, KA90]. Exercises [Rot92k]. Exercising [Ros90a]. Exhaust [Sch70a]. Exhibit [Ano95s, For80, Rot97c, Har96]. Exhibition [For80]. exhibitions [For80]. exiles [BR15]. Exist [Dan52, Mec22c, NA91m]. existence [Che20, Fre77b]. existential [Lin19, Mec17d, Mec23b, Pop19]. Existing [Pow47]. Exit [Mat12, Ram13a, BBL13, JP13, Mec12, Mec13a, Mec13b]. Exocet [Wal90c]. Expand [Cha68b, Ani15, McC88b, Riq15, Thr15, WF90]. Expanded [Kou75]. Expanding [BOW10, Liv68]. Expansion [Ano49-29, Coe98, Dea96, Woo96b]. expect [Dis79, Jr22b, Gol85b, Spe81]. Expectations [De 70a, Lan89a]. Expected [Foz62e, STS^{+}49, RS16]. eXpenditure [Wal81a]. Expenditures [Ano90p, Eps75c, Goo86, Leo64]. Expense [Bra47, Gol74]. Expensive [Ano18, Gre18, Hoa57, NA87h, Wil92b]. Experience [Bar73a, Fal70, Sto03, Why59, Al 11, Leb94, Bus85, Roc83a, Ack85, Lan76, Val73, Ano84h]. Experiences [Ano55x, Ano78o, Bor52b, Bru52, Cha52, Cot52, Cou52a, Dav52, Erd52, Fer52, Fri52, Gug52, Had52, Kas52, Ler52, Min52, Mon52, Oli52, Pei52, Pol52, Pri52, Sad52, Sch52, Wya52, SK84]. Experiment [Dvo65, Kow55, Sol70, Wat92, Wil72a, Ack88k, CP93a, Coc89, La 88c, NR79, Sel21, Kan77a, Ano56b]. Experimental [Ano69l, Bro56a, Fed70, Fra72, Led66a, Kat54c]. Experimentation [Ano77l, Gor96e]. Experiments [Cra49, Ras55, Zim75, Ano60i, Ano96u, Mat88]. Expert [Ano58e, Fin56, McR51, Nad66, Rat62, Cas76, Cro18, DM18, Dro19, Leg87, Mec16f, Mec23c, Men96b]. expertise [NR79, dF79d]. Experts [Ano60b, Ark02c, Coo76c, Mur54c, Nel74b, Ste75b, Hol76a, Mat79, MW91, WA93]. Explain [Ber93b, Sha10]. explained [Roc97b]. Explaining [Bro47c, Bro47c]. Explains [Bli70, McK23a, Mos93b]. Exploded [Bet46, Bet20, Wou78, Gre82b]. Explodes [Gom74]. Exploding [Gol69a, Lov00]. Exploiting [Eva87a]. Exploits [Ste72b]. Exploration [Ano50-37, Dys69b, Dys69c, Fol75, Gol66, Jac73e, Kat50a, OR70, Rab61i, Sag73, SJ68]. Explorations [Par88b, dRMT^{+}47]. Exploratory [Wyl56]. Explore [Rab62k]. Exploring [Ano12f, Lev77c, Sin70, Tur73, Lew72a]. Explosion [Ano49-34, Bet81b, Bro49, Gro49, Gui65, Ing65, Lal65a, MBD^{+}62, Orn67, Osb49a, Pen52, Rei59, Sei49, Ure49a, Woo62, Wyl56]. Explosions [All59a, Arn53, Bro60c, Bro60d, Dav79b, Don90, GLMS60, Joh60a, Joh60b, Lap48, Lap56a, Myr75, Spe58, Bar80g, Lon76a, Ste77c, Dav79b, Pre62]. Explosive [NA87a, SM90]. Explosives [Cla61, IS67, Rei50, Ste60a]. Export [Ano88e, Hev47, Alb94a, Brz89b]. exporters [Man77]. Exporting [Cha89b, The89a]. Exports [Ano87n, Ano89e, BC16]. Exposed [BD76, Sel00, Ric12]. Exposes [Mar49a, SM14]. Exposing [Pol93, Wur93b]. Exposition [Bri48b]. Exposure [Bio73, Bru57, Cum06, LS58, Mor59b, Roc97b, Alv80, Bey12b, DM18, Sal90b, Wil92a]. Exposures [Ato54b, Cla56, Fin90, Jac74b]. Express [Ano54-28, Kih47f]. Expropriation [Hag77]. Extend [Ano48q, Cle48c]. Extended [Ano48o, Ano48c, Tod78b, Mec23a]. Extending [Dam62]. extension [BT95, Eps95, Toy95]. Extermination [Tha63]. External [War49]. extinction [EE81b, SMĆ08, EE81a, Pos79, Har81]. extra [NA92j]. Extraterrestrial [Ehr71]. Extreme [Tsi82b, Wyl56]. Extremist [Sie05a]. Eye [Dow81f, Har92a, Bul03b, Ano95k, Isa89i, Rot95j, SC97, Sto96]. eye-in-the-sky [Isa89i]. Eyeball [Leb92, Leb92]. eyed [Ack89d, Ano90-31, Ano96k]. Eyes [Bre13, Dua78, Dys63b, Gom75a, Mos90, ZvH01, Kli90, Mec20d, MS77a, Rot93i]. eyeshade [Ano91-37]. Eyewitness [Ano52i, Mar91a, McC88a]. Ezra [Ano88j].

F [Abb91, Ack85, Ano78q, Ano90o, Bur48, Bur65, Bus85, Cra83b, Gra62, Gus83, Haw50, Jac73f, Kon67, LL90, LeR50a, Lyn69, Mar78a, Mar82b, Mar54f, Mat56, Pla57b, Pre62, Rab73d, She60, Sim54a, Sti83, Sto56b, Val86, Wri63a, Wyl56, You68a, Zac76, Isa99b]. F-22 [Isa99b]. F. [Ano47w]. Fabian [Bal79c]. Fable [Doo59, Dow80h, Dow82h]. fabric [Woo91]. Fabricating [Col93a]. Fabrice [Tho80d]. Face [Ast67, Bal56, Szi49b, Ark84b, Wat77, Rec89]. Facebook [Fie21b]. Faces [Ano46-106, Cla74, Dad67, Lap54d, Mes84b, Osb49a, Rab53g, Sal63b, Sha06, Ano92v, Sig87b, vH93]. Facial [Leo19]. Facilitator [Poc05]. Facilities [Coo81, Ano03i]. facility [C^{+}87b, Col79b, Abb04b]. Facing [Ano93l, Bro67a, Fin57a, Gol96a, Sau78, Soc07, Mec21e, Par78b, Dow84c, Fer57a]. Fact [Buc50, Cla68, Joh22, Ste75b, dF79h]. Factor [Net75, Rob77, Swe86, LH88, SAB^{+}16, Toy96]. Factories [All54a, Fre95]. Factors [Aug48c, Cro67, Mye87a, Sch68d, Sch47, Soh64]. Factory [Ano91-61, NA90k]. Facts [Ano54-30, Ano79i, Ano79m, BBSS50, Bye88, Car53, Com75h, Ell47, Kon67, Lov74b, Man89, Mit50, Ore58a, Rab65g, Rot97e, Som20, Szi49b, TvH83, Ano79f, Ano79g, Dom97, Fre76b, Kaf98, Pre96, Sin76, Siv76]. Faculties [Ass53]. Faculty [Ano96v, Ano96w]. fades [Lan90o]. Fahim [Ker68]. fail [Mos92, NA92e, Tin23, Woo93, Phe63]. fail-safe [NA92e, Phe63]. failed [Pat90a, Eri95]. Fails [Rab47d]. Failure [All06, Cir98, Eps75b, Kel03, Kih47f, Opp63b, Pai04c, Rab60f, Rya07, She60, Shi48a, Ano86-35, Ano94l, Fly98b, Gil18, Gol78b, Hen78a, Nei88, Sha93c, Wio79b, Jr22a]. failures [LaM79, Fis83]. Fair [Ano57g, Mil62, Mor53b, Tro79]. Fairfield [Pog54]. Fairies [Cha06a]. fairness [Bat88]. Fairy [Hib95b]. Faisal [Dro16b]. Faith [ACU^{+}54, Doo59, Har77a, Joh65c, Owe54, Sto55b, Tod80]. fake [Sto18c, TN90]. fakes [YSW21]. Falk [Lal67a, dF79b, Keu83, Wil84]. Falklands [Bar82c, Gol83b, Mot85]. Fall [Ano56h, Arn55, Cow65, Fer68, Hec70, Hol58, KS58, Lap55b, Lap55c, Lap55d, Lap55e, Lib55, Lib58, Mar71b, Neu58b, Oga58, Rab56b, Rab72a, RSW55, SM89b, WW58, Abr79, Boy95, Gut93b, Isa87f, Kun97, LGW92, McK25a, Pra22, Ano56c]. Fall-out [Ano56h, Arn55, Hol58, KS58, Lap55b, Lap55c, Lap55d, Lap55e, Lib55, Lib58, Neu58b, Oga58, Rab56b, RSW55, WW58, Ano56c]. Fallacies [Gou59]. Fallacy [Ack87a, Aft94b, HO88, Lac85, Yor69, Oel14]. fallen [Fie24, Isa90k]. Fallers [CMMF62]. Fallibility [Dum80]. Falling [Ben07, Ano98l, Ark94e]. Fallout [Ano61p, Ano61v, Ano62r, Ano70b, Ano93l, Aue06a, Bro70, Bro85c, Car82a, Cha63a, Foz61e, Foz62e, Gen59, Gou59, Ham63, Lap59d, Lap59e, Lap59f, MSDP80, MD64, Sch59c, Tam74, Wig60, Zel70, vBO62, Ack90f, Bar82c, BHM91, DM24b, Han97, Isa90f, Mar89d, SH94, SCHD20, WSG^{+}15, Gol88g]. Fallows [She82]. Falls [Est01, Mag88b, Sar83, TM96b, Tsi83h]. False [Coh75a, Har50a, RGK95, Sta93, NA90n, Pai82e, Wei13]. false-flag [Wei13]. Falsehoods [Bro85c]. falter [Gol93d]. Family [NKH03a, Wad74, Ano97j, Bro88a, KN14a, MB98, NA89b, NA90p, Wur93c, Cam54]. Famine [Cha69c, Fel86b]. Famous [Cha75a]. FAmS [Ano46-41]. fancies [Sin76]. fans [KA91]. Fantasies [GT06, Zuc82]. fantasy [Mec19h]. Far [Arn47, Cle48c, Kam05, Lev76, Rot04e, Rub74, Wei71, Wri04, Day77h, Fel77b, Fie19, Mak90, Soc15, Squ17b, KAT^{+}71, Sal79a, Wal50]. Far-Fetched [Lev76]. Far-Reaching [Wei71]. far-sighted [Fel77b]. Fare [Smi72a]. Farers [Smi72a]. Farewell [Bai92, Eis61, NA91l, Rab61e, Sie06b, Tru53, Uhl96, Day77f, Dar91]. Farida [Dro22h]. Farley [Fre84a]. Farm [Wea99, Bro88a, But84, Rot93l, GP92c]. Farmers [Rot92f]. Farms [Fre63]. Farrington [Ano46-39, Bro54]. Farther [Sch68f]. FAS [Kal82c, Ano47h, Ano47i, Ano48z, Ano49p, Ano49q, Ano50t, Ano52c, Ano53-42, Ano54l, Ano54m, Ano65g, Ano89p, BH53, Cou57, Fed63, Fed64, Kal82c, LW61, Nov56, Rob49, Sci48, Sch68b, Cou50, Spi49]. Fascell [BBG^{+}85]. Fashion [Aft00]. fashioned [Kel87]. Fashions [Rom90]. Fast [Abr74, Cal97, Cra83b, DeV95, Isa90g, Lov73, Wes00a, Dre89b, Gub15, Loe89, Nag15, Nuc18, Tak17, Zha15c, vHJ97b, Abb04b]. Fast-track [Cal97]. Faster [MB99, Smi58e, Smi58e]. Fat [IK01, Nor04, Rao77]. Fatal [Fra89, Pol88]. fatalism [Pet11]. Fate [Ano82e, Cha70b, McK74, Ano82n]. Fateful [Opp50, Rab56c, STC74, BPTB76a, BPTB76b]. Father [GR67, Rab05a, Ano89m]. fathers [Les93a]. Fatwas [Bla11]. Fault [Rot03d, Hun85a]. Faustian [Ano96n, Har76a]. Favor [Ing54a, KB88]. favored [Coh95b, Isa91b]. Favors [Ano46-91, McN48, Her88c]. fax [NA90d]. FBI [Ano91-29, Zei73]. Fear [Ark06, Ber64, Bid51, Bre93, CPV01, Eps75a, Gom00, Ing60b, Jac74b, Kih47f, MR07, Mue07b, Mue07a, Nor54, Pen77b, Sea68, SAB^{+}16, Sle84, Smi88, AC86, Ema21, Min86, Mül90b, Rot92f, Sny85, Wei15, Fra81, Ano81c]. Fearful [Rab60b]. Fearing [Mue07b]. Fears [Ler81, Mul69, Sal78b, Str05e, Afh88, Fau87, Fly97, Fra79, Gal91, Sub84, Wou78]. Feasability [Ano45p]. Feasibility [Fri54, Ost67]. Feasible [Aug48a, Ore59, Sha78b]. feast [Siv76]. features [NA91o]. February [Ano55y, Ano64p, Lei47, Tel57a, Cou57, Rab56d]. Feckless [Ano99n]. Federal [All55b, Can70, Dam61, Gor51a, Sto59b, Whe51, Ber84d, Row77, Ano51n, Ano53r, Ano53s, Ano65h, Ano67g, Ano78n, Con52, Dea83, Gro53a, Gro53b, Kay66, Mar65i, Pie66, Pow47, Rid51, Sap68, VR67, Wie62]. Federalists [Usb52]. Federation [Ano46-35, Ano46-108, Ano52y, Ano78m, Mey49a, ADH^{+}73, Ano45g, Ano45h, Ano46-42, Ano46-83, Ano46-89, Ano51i, GS51, Ano69k, Gro50, Hig48b]. Fedor [Leb91b]. Feed [Wad74]. Feedback [Bla61b, Ras75, Sch65a, Leg92a]. Feel [Ebe47, Ebe20, Leo19]. Feeling [Ano94m, Bur06, Kul88, You77c]. Feels [Ano48u]. feet [Kin78a, Ols16, Wes79a]. Feinberg [Ano78m]. Feld [Ano78l, Gab73, Tsi72a, Ano75a, Ano93p, BJH68, Fel69a, Ing82, Nat68a, WF68]. Felicitous [Sch87]. Felix [Doh57]. Fell [Lyn78]. Fellowship [Ano50b, Ano50o, Ano50x, HPL^{+}49]. Fellowships [Ano50a, Ano83q, OBC^{+}49, Rab49b, Szi49a, Szi63a]. Felony [Moo93f]. FEMA [Ano87w, DSL^{+}87, Pat90b, Rot93h]. fences [Mos92]. Ferenc [Mes85a]. ferment [Shu76]. Fermi [Ano68g, Ano94h, Cam54, Gro78b, Moo95d, Wil71, ASA55, And74b, And74a, Ano55k, Ano65f, Ano85h, Ano93-28, Hug57, Lan92a, Seg70a, Seg70b, Wei94a, vH01]. Fermilab [Wil78c]. Ferrari [Ano87n, Ano89e]. Ferreting [Dro24b]. Ferry [NA92k]. Fertility [Dje72b, Dje74]. Fertilizer [DPL75]. Feshbach [McC93]. Fetal [Sch59c, Tam69]. Fetched [Lev76]. Fetter [Ano90n]. Feudalism [Gla76c]. Fever [Mel00, Ark86h, Isa88a]. few [Bar80c, CB79, Man77, Rot95j]. Fewer [Ark86d, NA88k, Rep86, Che89, Gor95c, KK98b, Ple91]. Fialka [Hal93a]. Fiasco [Don90]. Fiction [Ano78n, Bar57, Bri87, Hor68, Hos87a, Hos87b, dF79h, Ack89a, Bri88, Fra89, Joh22, Klo86, Pei84, Rot92h, Sut16]. fictional [Woo97a]. Fictions [Hos87b, Rab65g, Fre76b, Hos87a, Rom90]. fidgeting [Rau96]. Field [BFW47, Cla56, Fel55, Gle51, Dow81e, Moo92l]. Fieldhouse [Ano90p, Han94]. Fields [Day88, Par73]. fieldwork [Dro25c]. Fifteen [Ano66f, Can63]. Fifth [Ahe91, Ano48-27, Ano49d, Ano49c, Lew49, Kal53, Mel53b, Mel53a, Say84]. Fifty [Moo95e, Pie54, Rob69, Was82, WB95, AO95a]. Fight [Hut53b, Lon78, Pos79, Smi63, Szi49a, Szi63a, Ano90-54, KSG15, Man15, McC90, McK23a, Pie22]. fighter [Ano92m, Mat98e]. fighters [Ste95]. Fighting [Ano87y, Sto18c, Woo96a, Bar80i, Pat81b, Thu79, Van88]. fights [Isa88c]. Figure [Gol09, Ano94k]. Figures [Maw78b, Zel70, Dro23c, Rot97e]. Figuring [Dro24c]. Fihn [Mec20a]. File [Les94, McM94a, Dow80c, Wel79]. files [GP92b, Mak95b, Zim89a]. Filling [NA92f, Sie06e]. Film [Ano81f, Ash81, Bar80m, Bar90b, Bar90c, Dow79b, Dow79e, Dow79d, Dow79a, Dow79c, Dow80e, Dow80g, Dow80f, Dow80c, Dow80a, Dow80b, Dow80d, Dow80h, Dow80j, Dow80i, Dow81g, Dow81f, Dow81i, Dow81b, Dow81e, Dow81c, Dow81d, Dow81h, Dow82d, Dow82a, Dow82c, Dow82f, Dow82b, Dow82e, Dow82g, Dow82h, Dow83b, Dow83c, Dow83e, Dow83a, Dow83f, Dow83d, Dow84d, Dow84b, Dow84a, Dow84c, Fla80b, Gue80, Hac81, HD82, Joh81a, Lad81, Mur84, Say84, Sch80c, Tho79a, Wea82, Ano84a, Dow77, Rot92h]. Films [Dow80e, Dow83f, Ash81, Rot83, BD79]. Final [EG24, Mye93b, Spi75, Zim81a, Ano98m, BM22, Hia80, Mec22d]. Finance [HK09]. finances [Rus92b]. Financial [WH16]. Financing [Gun20]. Finardi [Rez07b]. Find [Hir72, Bac80, Man77]. Finders [Kat61a]. Finding [Aue06b, Cle11, Jon13, WH00, Zag93c]. Findings [GNMS54, Sut83, Nor64, Shu78]. Findlay [Don90]. finds [Duf87]. Fine [Zel75, Ano92m, Ark95a, Ima93b, Moo95g]. Finer [Wri63a]. finesse [Bar92b]. Finger [Ark02d, Coo90]. fingerprint [Hib95f]. fingerprints [Hib94c]. Finish [CP99, Coo77, Ham90, Van88, DeV95, Sig88b]. Finland [Ste91a, Mie79]. Fire [Bon50, Bon51, Bon53a, Bro63a, Dje75, Ede04, Fly00d, Gro05, Han02, Isa02e, Joh81a, Kri85, Mar49f, Ros06b, Wei53c, Ano94o, Dro23m, Ede20, Hig77, Jac80a, Kla91, KM18, Ols16, Wea92, Weg96b, Wil80b, DeB81b]. Fireballs [Mue77]. Firebreaks [Dow84b]. Firecracker [Boy95, O’N94]. Fired [Fel83d]. firepower [Kla90]. Fires [Bro60b]. Fireworks [Wyl56, NA91e]. firing [Mac54a]. First [AT62, ABB^{+}50, And81, Ano46i, Ano46-43, Ano46-50, Ano47j, Ano55c, Ano57b, Ano63l, Ano63s, Ano83k, Bar79c, Bar66b, Ben76c, Ber73a, Bun82, Can63, Cha79a, Cle04a, Dow83d, Fed58, Fed76, Fel67d, Flo88, Gol78a, Got87, Gre73c, Isa91i, Ken79, Lal66i, Lan56b, Lap72, Len73, Lew62b, Lew68c, Mak03, Mil48a, Mil83b, Moo96c, Mor82a, Myr74, Nad98a, Nor63, Noy47, Pei84, Pri80, Rab57e, Rab59e, RF65, Sci46, She83a, Sti68, Sto76a, Szi55b, Tel48b, Wat74, You84b, ZG83, dTM93, Ano53o, Ano92r, Ark85h, Ark86e, Bli76, Car83c, CN91, CK20, Dea89, DG79, Dre89a, DuB62b, Eis62, EP86, Fel80e, GL90, Gro83, Gut93b, Hol20, KR12, La 79, LR79, Lei82b, Lie46, Mal79, May17, McC86b, Mil20, Mot83, Pat86c, Plo86, Rav83, Sak79, Sap84, Ste89c, SSR83]. first [Tho85a, TC83, Wei83b, Wil79, Ano55l, Kal89, Meh21, Pow86, Sap80b, Dro23j, Ram88b]. first-hand [McC86b]. First-Strike [Lap72]. Fiscal [Con65, Rei63, Tam71b, Ano87-30]. Fischer [Don86b, S.80, RK89a, Vuc69].

Fish [McL65, Roy67]. Fisheries [CP62, Pon67a]. Fisherman [Sch00e]. Fishing [Mar68, Roy67]. Fiskadoro [Bri86]. Fissban [Joh97c]. Fissile [AK04, GM09a, Rot97e, AH91g, Bel94, Hib95f, Rot77]. Fission [Alf74, Ano48p, Ano71d, Axe51g, Ben71a, FT76, Fel74b, GA74, Lov78b, Men48, Neu58a, O’K73, Bla11, Pau77, Iri56, Pri55]. Fissionable [ABB^{+}46, Pin57]. Fissioning [Bro75]. Fit [Wie47b, Ack89a, Dal92a, Dro21a]. fitness [Ano91-36]. fits [Isa96c]. Fitting [Bey12a]. Fitzgerald [Wri63a]. Fitzhugh [Bal79c]. Five [Ash60, Fra54, Kna63, Lal67e, Rab51b, Rab70e, Sel69a, Spi57a, Ste65a, Str06d, Ano85i, Ano89z, Ano90-62, Dow81d, Jr16, Dro21a, Men92a, Pai82i, Pat81c, GD85]. Five-year [Ash60, Pai82i]. Fix [CL06, Gri87b, Pod05, Pra04b, Dou24, Dun90, KB80, KR84, MM90, PWS^{+}96, vHWF^{+}81, BK81, O’B75]. fixed [Mye90]. fixer [Ano96l]. fixer-upper [Ano96l]. fixing [Moo18]. Fizzles [Isa99f, Her88a, Kub88]. Flack [Yeo75]. flag [Wei13]. flagging [Gol91a]. flam [Bel94]. Flanagan [Ano87l, Tsi89a, Tsi89a]. flank [NA88a]. Flap [Mos93a, Blu88, Bro88b, Eve88b, Rot96l]. flash [Alb94c, AG97]. flat [NA91b]. Flats [Ack01, NA89a, Aba89, Ack94, AZL90b, AGZ91]. Flavin [Roc85, Wou81]. Flaw [Lyn84, RB65]. Flawed [Kla95, Kra94, Kyl97]. Flaws [Lev85]. fleet [Spa95b]. Fleming [Gra62]. Flessati [Dow81f]. Fletcher [Mor61a, Wil68b, Bul08l]. Flexible [Vic78, Ark87e, Den82, Tra90]. Flexing [Mer90]. Flies [Sie05h, Ano95x, NA90d, Rot93r]. Flight [Bor58c, Dre84b, Tsi83a, Ram82]. Flights [Bla69b, Lei73, Lip61, Pat88]. flim [Bel94]. flim-flam [Bel94]. Flimsy [McM94a]. Flipped [Lea07]. Flipper [Ano94n, Rot96l]. flirts [AZ91]. Flock [Kre07]. Flood [Ber73b]. Flooding [Ada73]. floor [Sca14]. flop [KM90]. Florian [Har79b]. Florida [Cro65a, HK09]. Florin [Tho80d]. Flourishing [Mac89]. flout [Gol90a]. Flow [Ano78n, Dil66, Hak62]. Flower [PW76a, PW76b]. Flowers [Whi76a]. flunked [Rob90a]. Fluoridation [Alb64, Kor64, Mar64h]. flurry [Dun85, Kal85c]. Flux [Abb04b, Sch70b]. Fly [Ano69g, KM90, NA91g, TM96b]. Flyin [Ric96]. Flying [Ark84b, Ark98b, Hyn69, NA93e, Sag67, Spi61, Pra92, Sch76a]. Flynn [Sig85b]. Focus [AF85, Mar65a, Smi70a, Tam73a, Kan11, Kor19, Mec20h, SD86, YSW21, Dow81e]. Foes [Rot05b]. fog [Bur90c]. FOIA [Ano91-34]. Foiled [Ram88b]. Foiling [Ano96j]. Foley [Bas80]. Folk [Hil51]. Follies [Moo94i, Wei94b, AB89]. Follow [BV20, Ore62a, RS15]. Follow-up [Ore62a]. Following [Com75d, Rei78, Bro60b, Doy22]. Folly [Hur74, LP08, Pul04, Kis78, Pow98]. Food [Cal63, Cha68b, Cha69c, Com63a, Dow85, Lau76, LB78a, McG62, McP65, PGL91, Sax51, Sot67, Spu73, Sto68b, Tas51, Tod79, Aye85, Bis14, BL83c, DR14, Lou90, Mat18, Mor14, Ruf90, Wun82]. fool [Kre92b]. Foolishness [Mus75]. Footage [Dow83b]. Foothils [Ore60b]. Footnote [Oli89]. Force [Aug56b, Axe51k, Bar73a, Bec62, Bro63b, Fed58, Joh22, Lin84a, New64, Pal64a, Ros63, Wal55, Ano88k, BM99, Flo88, Hec12, Hec20, LS92, Mem79, Ano92m, Joh94a, Mar65d, McM54, Sto58, Bar66a, Rab53a, Roc81]. Forced [Mus01]. Forces [All59l, Ano84g, Ano91q, Ano99m, Ano99o, Ano99-27, Ano00o, Ano01b, Ano02i, Ark98c, Cal84a, Cra83a, Cra83b, Fin65, Kas69, Lal67g, NA90s, NA92l, NA96f, NA96d, NA97e, NA98g, NA98d, NA99a, NA99b, NA99c, NA99d, NA00a, NA00b, NA00e, NA00f, NAKH01a, NAKH01b, NAKH01c, NAKH01d, NAKH01e, NAKH02a, NAKH02b, NAKH02c, NAKH02d, NAKH02e, NK03a, NK03c, NK03d, NK04b, NK04c, NK05a, NK05b, NK05c, NK05e, NK05f, NC06, NK06b, NK06c, NK07a, NK07c, NK07d, NK07e, NK08a, NK08b, NK08c, NK08d, NK08e, NK09a, NK09d, NK09e, NK10d, NK10e, Roc81, Ano85x, Ano87-43, Ano92-31, Ano93-32, Ano98r, Ano99-29, Ano14a, Asp87, BH80, CAH84, DG07, For85, FLLW91a, Fre83, Gro86, Ken63c, Kok88a, KK19b, KK19d, KK19f, KK20d, KN11a, KN11b, KN11c, KN11d, KN11e, KN12a]. forces [KN12c, KN12d, KN13a, KN13c, KN13d, KN14c, KN14e, KN15a, KN15b, KN15c, KN15d, KN15e, KN16a, KN16b, KN16c, KN16d, KN17b, KN17c, KN17d, KN18a, KK18b, KND18, KN18c, KN18d, KK19c, KK20a, KK20b, KK20c, Lon85a, Lon95, NAC87, NA88s, NA88r, NA89n, NA90o, NA91n, NA92c, NA92i, NA93a, NA93b, NA93g, NA93n, NA93r, NA94d, NA94h, NA95d, NA95f, NA96a, NK10a, NK10c, PB^{+}01, Sen87, Sig87a, AM86b, Mor84a, Nor02a]. Ford [AM86b, Spi75, Sti83, Kas75, Pat83]. Fordham [Ano46-77]. Forearmed [Rab49d]. foreboding [Tur17]. Forecast [Bur47, Tru51, Wel51, Ehr84]. Forecasting [Han05]. Forecasts [Ack07, Hol73, Smi73]. foregone [Isa91f]. Foreign [Adv50, Ano46x, Ano46-29, Ano49-33, Ano66m, Ano67b, Ano69j, Ano89f, Bal84a, Bea47a, BBG^{+}85, Bla87a, Bow57, Dow88, Edg52, Fei60, Fei61, Fis61a, Fis83, Fit66, Ful69, Gol69b, Gru70a, Her83a, Hou72, How65b, Hum58, Isa04a, Kah50, Kin54, Kis60, Lan89b, Lon67, Mar75, Mas48a, Mer90, Mes84b, Moo95b, Mor67, Nel57, Nit57, Noy57, Rab67d, Sch69a, Sch71b, Sim54a, Spu72, Sta60b, Ste57a, Sti60, Tel47d, Tod78b, Wei54b, Wil59, Woo82, Wri63a, You66c, BL79, Bal85, Chu76, Cla81, Col84a, Dal83, Gor95a, LM80, Lyn88, Sch83, Sco08b, Mes84a]. Forensic [GA73]. forensics [Epp20]. forest [DeW83, Dro22i, TM96b]. Forestalling [Sim59c]. Forever [Ark00f, Sch05e, Ano97k, Ano98l, vHZC93]. Forewarned [Rab49d]. Foreword [Ber79, Day75j, Day77g, Gor51b, Hol86d, Lew71d, Pre62, Lyt79, Mar58a, Sei62, Sha54b, Tsi78a]. Forewords [Lak66]. Forget [Ano92s, Hir72, Kem97, Day77i, Tan88]. Forging [Som83a]. forgive [Osa77]. Forgotten [Lan98, Lar07, McG84, Kre98, Mag81, vCM89b, Zel92]. Forks [Day76d]. Form [Ano46-72, Ano55o, Fei71, Smi59b, Kis76b]. Formal [Bil85]. Formation [Lac57, O’K69]. formed [Hes78]. Former [Llu00a, Mec16b, Mec16c, Mec19c, Bor57b, Dro16a, Dro23a, Ric81, VA11]. formidable [Dah77]. Forms [Ano47h, Dan72, GMV^{+}24]. Formula [Gig60, Dre82a, Man77]. formulation [Sch87]. Forrestal [Joh51, Kap67, For49]. Forsaken [Kam05]. Forsberg [Ste95]. Fort [Ano54n, Ano55m, Mar54a]. Fortescue [Bal88]. Forties [Dow81e]. Fortifying [Isa99c]. Fortress [Isa99c]. fortune [Uri94b, Rab53f]. Forty [Ano89q, Ano90-56, Ano90-57, Fel85a, Gom82b, H.85, Moo97k, TvH17]. Forum [All55a, Ano83o, Bar63, Bou65, Cof65a, Cof65b, Con64, Dys63a, Etz63, Etz64, Fas66, Fru63, Fry65, Gal63, Gil63, Haf54, Hei64, Her67, Ist67, Jor63, Kol64, Lal64c, Lal64a, Lal66d, Lar63, Lar65, Led67, Leo64, Lil63, Mar65a, Mar66a, Mar64j, McN63c, New64, Nie66a, Orn67, Par66b, Pau63, Rab66c, Rav65, Rei67, RV64, Rub65, SS67, Sch65b, She64a, Sil64b, Sin63a, Smi63, Soh64, Spi66b, Sta63, Ste65a, Sti63, Sto65a, Sto65b, Wei64b, You63a, You64a, vL65]. Forward [Isa01b, Lan90g, Ano85j, Arb19, Bar10b, BHH11, Ekl77, Hei11, Kib10, Kor10, Lan10, Lan16, Mec15b, iM16, Nee79, Pic10, Puz93a, Ram10, Sep23, She16]. Fossil [Mus75, BW18, Fre77b, McK22]. Fossils [Fra72, Mag73]. Foster [Ano86i, Jac80b]. fostered [Ano85o]. fought [Col79c, Fot91b, GLH^{+}79, Hay79, Mil79]. Foul [Bai75a, Bai20, Gor96f]. Found [Foe07, Has68, Rab59o, TV71, Pow91a]. Foundation [Ano47o, Ano50-27, Ano50-43, Bar81i, Dub50a, Kat52b, Pet81, Pri47, Cou50, Tru50, Whe84b, FC88, Hin18, Ano47p, Ano50-38, Ano52l, Ano52s, CB95, DuB50b, Fro49, Gro52, Gro53c, May47a, Mey49b, Rab66a, Stu48, Sci49, Wat51, Spi75]. Foundations [Pol46, Whe84b]. Founder [Rab05a, Jr22a]. Founding [Aug56a]. Four [Ano55q, Ano71f, Ano79m, Ano87x, Fel72a, Her83a, Hum72, Lap72, Nuc95, Pan54, Sco72, Shu87, SM81, Sør95, Van60, VFIS48, Way51, Gar77a, Hen96, Kis82, Wei82b, Ano86u, Bin54, Dan54b, Fra54, Tho86a]. four-step [Hen96]. Fourth [Ano48-28, Ano48-29, Ing59b, RR68, Whi64, Pep15, Ano59d, Jac71b, Kal83a]. Fourth-Country [Ing59b]. Fox [Sco74]. fracking [Dro16a]. fraction [Mye90]. Fractured [Kin67, Squ17b]. Fragile [AC85, Eva99, EG24, McK25a]. Fragments [Moo00a, Pal00]. fragrant [Day77o]. frailty [Blo85]. frame [Rof19]. Framework [Kal60, Nel74a]. France [Pat81a, Wal83b, Ano90-58, Ano98p, Aro60, BPTB76a, Bes94, Bre69b, Dil57, Fer68, FC86, Gil50, Jol46, Kow48, Lap76, Nor56, NA90g, Per55, Sch13a, Sig85b, Sim46b, Ter04, Wal90c, Wea79, Zim76, dJ64]. Frances [Diw79, Dow82g, Ano14c]. Franchise [Ano54o]. Francis [All59l, Dow82f, Gre71, Pla57b, Pon67a, Rab56h, Wol55, Jr22c, Pep15]. Francisco [Fis81, Gei81, Mac81b, Mar64d]. Franck [Rab64b, Vil75]. Franco [Ne’81]. Francois [PF75, Rab73d, Sig85b]. frangible [Bar80d]. Frank [Ano85d, Day75c, For80, Rab70a, Siv76, Wyl56, dF79e, Ano15b]. Frankel [Cal82, Pla69]. Frankenstein [Bar69]. Frankensteins [All16]. Franklin [Ano86i, Bri86, Dow81a, Mar78a, Aur71d, Bri89, Gou75a]. Franklyn [Roc81, You66c]. Franks [Ark95b]. Frans [Rot93a]. frantic [Kar88c]. Franz [Mul66, Moo94b]. Frappe [Pal64a]. Fraser [Boy87]. Frauenglass [Ein53]. Fred [Ano85c, Hig66, Met83, Pat79a, Pol91b, Pos79, Rod85, Win90]. Frederic [Gol95d, Sei69, Ull63, Rab58b]. Frederick [Dow80h, Ham61a, Ras55]. Frederik [McC88a]. Free [Ano52j, Ano74b, Ano78p, Bar66b, Ber78b, Bur50, Das68, Gol93d, Jan73, Moo93b, Mor47b, PMRP50, Pie68, Pol50b, Ric53, Smy50a, Bul09a, Van90b, Wal08, vB67, And85, Ano89-28, Ano90-39, Ano90-58, Ano95i, Ano96o, Bou11, Don85, Fry85, Gar78a, Gri85, Mye90, ML82, NA93e, RF21, Rot94g, RP82, She94, SB23, TD89, Zag94f, Zei85, Day96]. free-for-all [Ano90-58]. Free-Vanunu [Day96]. freebies [Ano95j, NA89b]. Freed [Hig66]. Freedman [Roc83a, Mos81a, Wal87a]. Freedom [Ame53, Ano55p, Ano78p, Ber56, Boe57a, Boe57b, Bok49, Che50, Cus49, DS47a, Dar57, Ded57, Din69, Don57, Dys07, Git74, Gla56a, Gre67a, Har60b, HUC48, Hux49, Inf62, Kom54, Lev77a, Mas48b, Mor48b, Mul60, Mun65, PD49, Pol46, Pol50a, Rab46d, Rab68c, Sch68c, Shi47d, Shi57, Ure48c, Wie48, Zag94e, vL48b, Gar85, GS81, Har89b, Nat85a, Pai82h, Sak78, Sto59b, Sho92b, Sto55b, Ash48]. Freely [Apt64]. Freeman [Ano93n, Gom84a, Rea80, Rom89, Tuc97]. freewheeling [Cha89b, The89a]. Freeze [Ano83g, AuC84, Dot65, Fer88, Ing84, Kal82b, Lal64j, Led71a, Pai83e, Swe89a, ASB^{+}84, BL83b, Dow82i, Eps82a, Fel82e, Fes76, Gle83, Hal83, Isa82b, Kep82, LC88, McC81, Min88, Pai82f, Pai83d, Wil82a, Mor85a]. Freezing [Lla70, Ste95]. Freis [Ols62b]. Fremd [Dow81g]. French [For80, Han94, Kol84, Ahm03a, Alb84, Ano45m, Ano48v, Ano49r, Ano52g, Ano67j, Ano68h, Ano99o, Bar92b, Bon57, Bon58a, Bon58b, BBL13, Bru52, Cha52, Cot52, Cou52a, Dan90b, Dan91, Dav94, FC84, FC88, Fri52, Gal81b, Gol62a, Gol62b, Had52, JCKG^{+}50, Kas52, KM89, KK19c, KKJ23a, Ler52, Mar53c, Mec13a, Mil84, Mon52, NP80, NA87c, NA92e, NA93g, NBF94, NA96a, NA99b, NA00b, NAKH01c, NK05b, NK08b, Rin52, Rot94d, Sad52, Sch13b, Sch52, Sen65b, Sen67, Wya52, dR49]. Frenkel [Hol86a]. fresh [Has88, Pai83e]. fret [Rot88]. Freudberg [Dow81b]. Friction [Str05d]. Friedensgestaltung [RK87a]. Friedman [Mar98c, Mar98c]. Friedrich [Bar79b, For80]. Friend [Sen65b, Str78, Sha96]. Friendly [McC93, Ros06b, Ban22, Flo90]. Friends [Ano93r, Ano97j, Bat52, Ein49, Fel79b, Lat91, Rab52b, Rab59l, She05b, Ark97a, Gol96b, McM94b, Mil89a, Mec21c, Osa77, Woh78]. friendship [Wil78c]. frigates [Ano95n]. Frightening [MG08, Bro85i]. Frisch [BGR61, Fel79o, Bro47c]. Fritz [Asi62a, She61, Pat82a]. Fromm [Las62]. Front [Gra93, AH91m, Ano98m, Ano11d, NA92e, Sti18, Swe79a, Zam92a, Buk97a, Nel90]. Frontera [Fly02a]. Frontier [Ano86l, Ano98g, Deu83b, Gre08, Ken61a, Ano98m, Ban22, Mec22d, Tho80d]. Frontiers [Lal66f, Liv68, Tuc72, Ano96n]. Frozen [Rab57f, Ret99b, Wil80b]. FRS [BM99]. frugality [Hen78b]. Fruit [Ano69g, NA91i]. Fruits [Edm74]. Frustrated [Lan90k]. fry [Joh94f]. Frye [Sto58]. frying [Wea92]. fu [Fra81]. Fuchs [Ano50r, Che52, HM90b, MHM90]. Fuel [AH03b, Alv02, Axt73, Buk02, CL06, CD77, Hol74, Joh59, LK10, O’B82, STC74, Sue61, AF88, BWA85, Bas86, BW94, BW18, CJLT24, Dah78, FvHA86, Hei11, Jas24, Kan11, KKL15, Kaw78, Mos80, Nel12, Pan23, Rad97, RL14, RW80]. Fuel-Cycle [CL06]. fuel/weapons [FvHA86]. Fueling [Kla91, Wes00a]. Fuels [NAD^{+}47, Ant76, Fly97, Mac22, McK22, Mus79]. Fujimori [Rad92a].

Fukushima [Cha11, Coo11, DM24b, Jr16, Fel14, Fri11, FK12, FD23, Gla12, Ito14, Jor11, Kat16, Lym11, NF13, NGG11, Per11, Per13, Ram13a, Sch16, Set16, Sim16, Suz11, Tak11, vH11]. Fukuyama [Jr22c]. Fulbright [Cha68a, Gol67a, Sch71a]. fulfilling [Fre76a]. fulfillment [Sim11]. Full [Isa00c, Jac74b, Pri77, Ste91a, Dav92, Moo95q, War14]. Fuller [Lyn78, Boe76, Day75b]. Fulton [Ano46-107]. Fumbling [Ber70b]. Fun [Ano94k, Ano95j, Rot04c]. Function [Dre74]. Functions [Ano53t, MEA^{+}47, Opp47b, dRVT^{+}47, Zlo47]. Fund [Fra83, Ano82s, Ano87-36, Ano87-37, Ano88r, Ano89x, Mec21d]. Fundación [Ano91u]. Fundamental [Bet75a]. fundamentalist [ST22]. funded [Har89a]. Funding [Ano78n, AL79b, Jac74b, Ada86c, BV20, MRH16, NK86b, Rab60e]. Funds [Ano53r, Ano53s, Ano58a, Ano67g, Pie66, Gol86b]. Funereal [Aur71a]. funk [Isa91c]. Funny [Mus99a, Bro85i]. Furniss [Mul66]. furniture [Wur97]. Furor [AH03b]. Further [Ano90z, Bec64, Fed63, Fei81c, Gil69a, GLMS60, Gus82, Lei68, Len60, Mor62a, WZSA73, Wyl56, Gar88a, vH76b, Lew70f]. Furthermore [Fai83, JP82, Kol81, Mar83a, Mar83d]. fuse [Fel81i]. Fusion [Ano71d, Dro24f, Gom78, GR62, Hol78, Jac71b, Kat92, Pos71, Pos24, Rus50, Dro24a, Dro24b, GG11a, Gol24, Jas24, Lov89, Mak24, Rot97i, SB80, Dro24d]. Futility [Jer77, Whi78a]. Futterman [Mea62]. Future [And58, Ano45i, Ano59a, Ano60k, Ano69o, Ano79a, Aro60, Bac48, Bor73, Buc60, Cla74, Coc50, Coc55, Jr16, DuB69b, Dub56a, Dys60, Dys75, Eds58, Fed72, Fri64, Fri70b, Gol69b, Han05, Har56, Hoa72, Hux59b, Kap62, Kno60, Lal65c, Lal67h, Lin73, LK10, Mad79, Mar98c, Mar71a, McE52, Mea62, Mei52, Moc56, Moo93a, Nix10, Ore58a, PB68, Par86, Pat80c, Pot70, RK93, Rub65, Sea68, Sea70b, Sem59, Sem64, She79, Shi67, Sho92b, SS73, Spe68, Spi75, Spr69, Swe86, Tel51a, Tor61, Wou81, You66b, All15a, All16, APR17, Bac77a, Bar85a, Bar92b, Bel20a, Bow89, BHH11, Bro82, Bro84a, Bro85d, Cla16, Dal92a, Dav76b, Der85, Fla87, Flo85, Gre83a, Hin18, Hol16, Isa92c, Jac15, Kop15, Leo19, LvH18, Mac22, May17, Mec20e, Mec20g, Men97, Nye13, OW15]. future [Rav84, Ric98a, RK87a, SSK10, Sca11, Seg13, Sim84d, Suz11, WH11, Wal21a, WKK23, WL21, Bas80, Fer57a, Wou81, Ale69, Ano85c, Ano87l, Bla85a, Fri70b, Gan83a, Gro78a, Jac80b, Kat54f, Mar78a, Mit54, Rab64a, Ran85, Sto55a, Tod80, Wun82, Pat82a]. Futures [FT76, Man85, Rea80, Soc15, Sør95, Bas80, Pat81a]. Futurology [Ano92y]. Fuzzy [Pai85f]. FY [Isa80]. Fyodorov [Gui77].

G [Ano57f, Ano78o, Ano85d, Ano98g, BD79, Cav61, Day75b, Deu83a, Gan83b, Gol95a, Gre54, Gre71, Gre82b, Gre87, Her90b, Kat52b, Lap55a, Lap61a, Leb94, Lyn78, Mul65, Nie63, Per87b, PF75, Ras55, Rep85a, Saw80, Sea72b, Sin67a, Smi58f, Sto56b, Tsi72a, Why62, Wol55, You56, dF79h, HB10, Joh03]. G-8 [HB10]. G-Man [Joh03]. Gabriel [Mer90]. Gabriele [Len21]. GAC [Lap59g, Lap60a]. Gaddis [Gar97a, She83b]. gadgets [Moo95d]. Gaffney [Isa88c]. Gagarin [Lip61]. Gaia [Ano93n]. Gaidar [Gut93d]. Gail [Wal81a]. Gain [Ano90-27, Pod01]. Gains [Jac74b]. Gal [Dow80i]. Galactic [Haf67]. Galápagos [Bri86]. Galbraith [Eis67]. Galilei [Sha54b]. Galileo [Cal57, Dub57a, Sha54b, Sha54b]. Gall [Sta63]. Gallagher [Pol93]. Gallery [Gen72]. Gallois [Sei62, Sta63]. gallon [Bro80b]. Galluping [Sag50]. gambits [Sto85, Swe85]. Gamble [Wal84, Fel78e]. Gamblers [Hoo02, Hoo20]. Gambling [Sha81c]. Game [Ano78n, Aue03c, EB73a, Hal93a, Kar80, Kin86, Mar65e, Moo02b, Pol73, Rab51h, Woo04, Ano82l, Ano91-40, Ano91-42, Ano97k, Ben83, BL83a, Bro85e, CM91, Coo79, DM82, Day76g, Dor85a, Eri88a, Eri88b, Fel76b, Gal78a, Gou78a, Ise92, Kal83c, Kun23, Lei82b, Nol90, Pai85b, Pos87, Rot98i, Spe89b, Sta80, Dow82h, Sto17c, Ada87, Dow84b, Mes83b, HP80]. game-changing [Kun23]. Games [BB79, Hyd75, Pil82, Rot04c, Sch06b, Ano91-35, Ano94k, Ano94m, Gai92, Hal93a, Isa92j, Men92b]. Gamesmanship [Hum72]. Gamma [Bla02, Sal90a, Moo95a]. Gamow [Wil71]. Gander [Rab67i]. Gandhi [GD85]. gang [Gar77a, Dow79c]. Ganguly [Das96]. Gansler [Nas82]. GAO [Flo88]. Gap [Bla69a, Bru72, Her67, Lop68, Rab67c, Ric03c, Sie06c, Sot67, Tam71a, Zam95d, Ano92r, Fly97, Gar84b, Kaz14, Mal14, Nie14, Odh79, Sco84a, Sin14, Slo12]. Gaps [Lal67d, Ger17, Rom83, Tsi83g]. Gar [Hig66]. garage [Ste83b, Sha84a]. garbage [Sal90a]. garble [Ark95a]. Garcia [Fel82g, Bar67b]. Gardener [Sch00e]. gardens [Ano80o]. Gardner [Hyd75]. Garey [Dow80e]. Gareyev [Hol88]. Garfield [Ano78o]. Garfinkle [Mor85a]. Garlin [Sto18c]. Garrett [Hul78, Lew72a, Lew73a, Ano12d]. Garrison [GNMS54]. Garry [Kar80]. Garthoff [Fre88, Gal86, Uhl95a]. Gartmann [Ano56d]. Garvey [Nel74a]. Garwin [Mar78a]. Gary [Bar90b, Dow84a, Mor87c, Nor64]. Gas [Ano53-46, Ano78n, Gol70a, Jac72a, Mar65g, Mes71, Mes72, Mil71, Sie06d, Wil73c, AH91i, AH92b, BC16, Isa83a, Klu22, Lev12, Pay84, Ram16, Tan88, Wol78]. gas-cooled [Ram16]. gases [Ano60i, Spi88]. gasps [Isa88f]. gastronomy [Ano98n]. Gates [Ano92x]. Gateway [Isa94i]. gathering [Kib12, Nay12, Squ12]. GATT [Rei83]. Gaulle [Gal63]. Gave [HM90a, AM21, Isa89l]. Gay [Ano95s, Har96, Rot97c]. gays [Rot93m]. GCD [Wie68]. GE [Zhe90c]. Gears [Wur99]. Geburtstag [For80]. Gedächtnisausstellung [For80]. Geis [Tho80d]. Gelb [dF79d]. Gellhorn [Ano52-28, Rus51a]. Gellman [Iri86]. gemeinverständlich [Thi46]. Gen [Sch06c, Ber23]. gender [Kaz14, Mal14, Nie14, Sin14]. Gendering [Tic94]. Gene [Eav18, Tay87, Che88]. General [Ano47g, Ano47q, Ano50q, Ano52a, Ano57j, Ber03, For80, Gol88b, Har60a, Lap69, Ler80a, Noy47, Ros67, Was62, Whi53, Ano79k, Ekl77, Isa89l, Mec21g, Tor19, Wil79, Ano45i, Ano46-31, Ano85t, Ano85w, Ark87c, Bra49, Dew52, DP12, Fry57b, GNMS54, Gro49, Jac67, Mec21c, Nor02b, Nor03, Rab48f]. Generalist [Zim75]. generals [Cra96, Ler80a, Lop90, WN79, Kib01]. generate [AGB22]. Generation [Eli71, Fen91, KK02, Lei07, Rab67e, Ano15f, Hic22, Kap91a, Mag16, NA94e, Nye77, Rot93e]. generations [Tay86]. Generis [Pop52]. generous [Ano97k]. Genes [Mul47]. Genesis [Dow81a, Hei70, Hen71a, Sch07a, Tho80g, Sal87]. Genetic [Car55, Cha69c, Che88, Cla56, Cro58, Dan72, DeN76b, Dje72a, Fre72b, GC83, Hoa72, Jai73, Koc77, Lin08, McD74b, Mul55b, Mul55a, Mul68, Nag71, Pau62, Pau63, Pau70, Ram72, Rus56b, RGR63, Wes55, Ano77l, Cav82, Mec20f, Ram72, O’B75]. Geneticists [Dun49, Sax49, SSW^{+}55, Wri49]. Genetics [Ano49v, Ano55n, Dob52, Jac68, Jen70, Led66a, Lew70f, Mul48, Rab49e, Rab55b, Sax49, Lan56a]. Geneva [Ano55l, Ano55q, Ano59i, Ano60r, Ano79h, Ano96i, Bar80c, Bel72, Boh55, Chu60, Eat62, Fol60, Isa85e, Joh96c, Lal67e, Lap55f, LN83, May47b, Mes72, Nor56, Ore60a, Rab55a, Rab55b, Rab60f, Ran55b, SW56, SW59, Smi55, Tur60, Wei55a, Wei59a, Wei55c, You61, Zam96b]. Genie [Moo97b, Moo96j]. Genius [Doh58, LS94, Pow88, Cha89a, Low93]. Gennadi [Gut93b]. Genocide [Cas74, Jac74b, Tho80g]. genomic [Hil12]. Gentle [Com75b, Loc89a]. Gentlemen [Coh53b]. Geodynamics [Dra73]. Geoengineering [Ham14, Rob08, Ano14g, Ano15a, Sch19]. Geoff [Pat81a]. geographical [KKS22]. Geography [Rab72c, Rob90a]. geologic [BM22, Tra24]. Geological [Ano59f, EG24]. Geology [Cra70a, Rub70, Smi72c, SS73]. Geophysical [Fed59c, Sim57a, Tro58, Wee73]. Geopolitics [Ray81, Pal16]. George [Ad 90, Ano78n, Ano85a, Bou98b, Cle74, Gus83, Hor77, Kat52a, Kre93, Lyn69, Mar98c, McK74, Mul73b, Nor64, Ost72, Rea80, Sch68f, Sim54a, Tho80d, Vuc69, WW62, Wig62a, Wil71, Wur68, Ano86v, Ano97p, Ano13b, Ano13a, Isa89j, Moo91g, Rat83, Rot92g, Rot04d, Bul10, Wol58a]. Georges [Ano60f, Ano85h]. Georgia [Uri94a, WM78]. geoscience [NGG11]. geostrategic [BC16]. Geothermal [Rex71]. Gerald [Abb85, Ano78n, Ano78m, Ano85b, Bar66a, Bas80, For80, Har92a, Nel74a, Rei63, Sim87a, Sim85, Sne57, Tay91]. Gerard [Sim62]. Gerber [Gra93]. Gerd [Swe86]. Gerhard [Moo99e]. Germ [Ano52c, McE88]. German [Mul66, Rab68g, RK87a, Sel63, Swe86, Swe93, Wei48, Ass63, Ano65e, Ano68l, Ano75g, Ano91-30, Ano93-31, BS47, Bet93, Bir89, BBF^{+}57, Brz89b, Bur62b, GP92a, Gou47a, Gou47b, Gya90, Hib95c, Hun14, Jan15, Kru81, Lal65b, LRD^{+}91, Mec12, Mor47a, Mul69, NP80, NA90p, Pow93, Rab68g, RK88, RF65, Smi67, Thi46, Thr15, Wal90b, vL47, vL48a, vL48c, Bar58a]. Germans [Gom75c, Hib96a, Thr92]. Germany [Ano55p, Ano89g, Ein48a, RK89a, RK93, The89a, Ack09, Ano46c, Cha89b, Dea83, Fel14, Fri21, Gou46, Gou48, Hei47, Hib01, JvH09, Kün96, Lal65e, Mat12, Mei20, Mez12, Mor50a, Mue01, Mül90b, Pel64, RK87b, Sch04a, SS89, SOE21, Som84d, Wes04, vW57, vW59, Spe68, Wal83b, Mul72]. germplasm [Wil77a]. Germs [Cho01]. Gernot [Lyt79]. Gertrud [Gab73, Fel81m]. Gervasi [Jon87]. Geschichte [Thi46, Wei48]. Get [Ahm04, Alb06, Ano97k, Ano97m, Cle04a, Hit01, Luc90a, Mac67, Ram88b, Sch06b, Ano95t, Eps94, Flo89a, Gei17, Hic22, Ing76, NA90k, Obu89, Sha90, Sok20, Zel76a]. Gets [Fer99, Ano91-34, Arn96a, Rot94e]. Getting [Ano95j, Ano11g, Bar10a, BKV13, CP98, Cum15, DE95, Gor96a, Jac71a, KR11, Lev84a, Loc95, LH99, LC94, MRH16, Sig87c, Bul09e, Wal07, Ano90-33]. Gettleman [Sch68f]. GeV [Ano68c]. Ghastly [Mat88]. Ghat [Lar62a]. Ghoshroy [Ano10a]. ghost [Ano97q, Dro23j, Rof97]. Giant [Her90b]. Giants [Com74, Mar79b]. Gift [Ano98h, MH89b, Moo94e]. Gifted [Har56]. Gilchrist [Sto18c]. Gilmore [Lyn69, Tho80g]. Gilpin [Mes64b, Wig62b]. Gimme [Ano03h]. Ginzburg [Sto59b]. Giorgio [Sha54b]. Giovannitti [Hig66]. Gipper [Cai89]. girl [Fri20]. Girls [MM01]. GITSM [Vos88]. Give [BFM^{+}49, CFM^{+}10, EBU^{+}52, Eps94, Isa92m, Pow90b, Wil89, Ano92r, Lat91, Man77, Row77]. giveaways [Lum96]. Given [Ano49-33, Car03]. gives [Abr79, Day96]. Giving [Ano94o, Bel02, Hol94d, Kor92, Wei93, Che52]. Glacier [Dro20a]. glaciers [Hus24, Mer24]. glaciologist [Dro24c]. Glamour [Smi75a]. glance [Ano90-59, Men91b]. Glare [Ros90a]. Glaring [Ger17]. Glasnost [Ano92x, Ano93y, Ano93q, Kre90, Leb91b, Tel91, ML88, Wol88, Fra87, Ano90o]. Glass [Ano55i, Cum99a, GMV^{+}24, Moo95q, Shu78, Squ17b, NA93j, Rob56]. glasses [Dor85b]. Glasstone [Ano78f, Mol62a]. GLCMs [Flo89b]. Glebov [Ano90o]. Glen [Dow80a]. Glenn [Lon82, Tho80d, Wol63, Moo99e]. glick [Coh87, Bie68]. Glimcher [Dow80h]. Glimpse [Vol74]. Glimpses [For80]. Global [Ano88e, Ano89w, Ano90o, Bar85a, Bas80, Ber73c, Bor70, Cha74, Cox86, Dad71, EE85, FR97a, Fle70, Flo85, Fre76a, Gan82, Gle94, Gol95d, Gol99, GM09b, HIKWB10, Hol75, Ind81, Jac73g, Kap81a, KN13b, Lan90i, Lap55c, Leg92b, LG15, Loe02, LSP95a, Mas67, MF10, MM90, Moo97b, Mou73, NA97b, NA00c, NK02, NK06a, NK10b, Pol09b, Pre96, Rab98, RR68, Sch90, Sch78a, SL58, Som84c, Sto72, Tod79, Van22, Wea11, WH72, Wio81, Ano90-34, AB18, Ark86a, Bol88a, Bud22, BKV13, CG12, DP20, Fou84, Han17, Kea24, KSG15, Kul17, La 88a, Mac91, Mar87b, Mec23c, Mos80, Nor11, Nuc18, PFW18, Raj17, RK19, Rav84, RTB^{+}19, Sco08b, Sep23, She14, Sim84d, Sto17c, War78, Wea92, Fri89, Mye87a, Nad98a, Smi90a]. Globalism [Nad98a]. Glomar [Ano92q]. gloom [HF80]. Gloria [Win84b]. Glorified [Whi77]. Glory [Jos02, Les93a]. Glossary [Ano55n, Ano58f, Ano59f, Ano61l, Ano75i, Ano86-33, Ano86-34, Pri75c, Ano88h]. Glow [Ano92q]. glowing [Ark97e]. glows [Ark93b]. glut [NA93f]. Gluttons [Zim77]. Glynn [Ise92]. Gnats [Wil75g]. GNP [Dar71, Ois72]. Go [bZAAFC08, Ano59a, Ano90-28, Ano94k, Ano96k, Com75b, Eps75c, Fel76d, Isa00f, Joh96a, Ken60a, Lil49d, Lon67, Lon73, MZM01, Mue01, Nan71, PS04, Roy76, Ram64, Ros75, Rub74, Tay89, Ano90-32, Ano98l, Kre92b, NA89j, Rat92, Rev81, Woo98c]. Go-banken-sama [Joh96a]. go-it-alone [Rat92]. Goal [Lew67a, SG62b, Eps93, Lin86]. Goals [Hek60, Kel64, Par71, Wis55, Lur77, Squ12]. goat [Ano95j]. God [CO14, Dow80f, Moo94e, Pop52, Sma76a, Mar69, Fly98a, Rap65a]. Goddess [Doh58]. Godfather [Nor06]. Gods [dF79h]. Goebbels [Ano48k, Zag95c]. Goes [Gol99, Isa03g, Kev07, Ros90a, AT88, Ano92u, Ano97i, Ark97f, Deu84b, Eve89, FP17, Isa84b, Isa03c, NA93i]. goeth [NA89c]. Gofman [Cow71, Pau70, Rad83, TG70, TB70, Wil92a]. Going [Ano94k, Ano98o, Ano97t, Ark84c, Gub15, Nag15, NA93p, Say84, Bul02c, Zha15c, Ano98m, Ano14b, Bro80b, Dro21d, Kim20, NA89s, NA93t, Bro80b]. Gold [Ust93c, Wal81a, Zag95b, AH92a, Ano91-37, Ano98o, Kre92b]. Goldberg [GJ67]. Goldblat [Ano88e, Don86b]. Golden [Her90b, Ken77, Isa89k, Isa90l, Pow90b, Her88b]. Goldman [Mei90, Mil88]. Goldschmidt [Fel83a]. Goldsmith [Ano51o, Ano52-28, Ano49w, Dow80f, dF79f]. Goldwater [Ano86v]. Golem [Rap65a]. golf [Ano94j]. Golovine [Sch63c]. Golstein [Fel92]. Golub [Ano60f]. Gompert [Mar78a]. Gondwanaland [Cra70a]. Gone [Arn47, Col03, Coo75a, Lor00b, Bro80b, NA89c, Rot98c, Swe89b]. Good [Ano90-31, Ano92y, Aug50, Bru88, Cou71, Fel77d, Guz52, Kas65, LB78b, Man64, Mec18a, Mos52, NA89f, Pra05b, RMK65, Rot96e, Rot03b, Rot04e, Sam18, Spe85b, Sto78b, Val52, War88, Wes74b, Whe84b, Whi53, Woo98f, dO52, vA73, Ano90-47, Ano94j, Ano96n, Ano97i, Dea89, EG89, Isa97a, Kis79a, Kra15, Kul88, Ram88a, Rev81, Sch11, BB63]. Goodbye [Ano97n, Iva02, Meh21, SOE21]. Goodchild [Ano05b]. Goodell [Dro23j, dF79e]. Goodin [Ano85b]. goods [Kis93d, Rot94e]. Goodwill [FP90b, McC84]. Google [Squ16]. Goonatilake [Eas85]. Goose [Rab67i]. Gorbachev [Ano85k, Ano86t, Ark89b, Bro88a, Bro85a, Bro86b, Bur88, Cra87, Dea87, Gol85b, Hof88, Jac92, Mcc88c, Smi89, vH87, Tum87, Uhl97b, Mil88]. Gorby [Ano92u]. Gordon [Ano50-41, Ano78n, Smi54c, Yud83, Mur58, RS52]. Gorgon [De 70b]. Gorky [Eva87b, Sak83]. Gorleben [Pat79g]. Goro [Dow81i]. gossip [Ano95n]. Got [Moo03a, Coh95a, Fri89, Kar17, NA93l, OB78]. Gottemoeller [Ano11g]. Göttingen [Got83]. Gottlieb [Uhl97a]. Götz [Ano10a]. Goudsmit [LRD^{+}91, Wal90b]. Gould [Mei90]. Goure [Fel73a, Mac62]. gout [Gra81a]. governance [JFS20, MvE13, Sep23]. Governing [Ano47q, MGZH07, MA17]. Government [All55b, Ano46-66, Ano50-36, Ano52k, Ano53l, Ano55w, Ano61r, App46, Bea67, Bec55, Ben71b, Bha71, Bro47e, Bro47f, Con48, Con49, Den47, Doo64, Dre74, DuB66, Ell56, Fin56, Fox46, GR62, Kal89, Kor92, Kre11, Lal65d, Lis46b, Lon86a, Mar85a, Men68, Moo51, MW47, Nea02, Nie49, Par66c, Pow47, Rab47f, Rus58b, Sut68a, SG69b, Tay47, Ure48a, Usb47, UH56, VFIS63, VR67, Wal51, Wei85b, Wis55, Wri63b, Zei51, Dan54a, DeW79, Dor85a, Dou15, Dro23c, Ger20, Ols16, Rus20, Sam18, Shi47d, Zag95c, Ano53o, Ano53y, Ano58c, Con52, Joh58, Lar50, Orr56a, Pow47, Rid51, Whi51b]. Government/Press [Wei85b]. Governmental [Rod61, Wol68]. Governments [Abr79, Bam85, Int50, Pau57]. governor [SD88, Ano48r, Ano48-30, Ano53v, Lap54a, McM48a]. Gowing [You76a]. Grab [Web02, Lee91b]. graced [Sag83]. Grade [Kup06, Sin05b, Rot77]. Graders [Say84]. Grading [Fre91b]. Gradual [Etz63]. Graduate [Abe74]. Graduated [Osg60, Sha61]. Graeme [Gri83a]. Graham [Deu83b, Wal85a, Ano10b, Mec22b, Soy89]. grain [BB79]. Granaries [Bor73]. Grand [Bar88, Cah76, Mul66, Ric71, Day76d, Mos67]. grandchild [NA89i]. grandparent [TS22]. granny [Ano97n]. Grant [Ano61m, Ano87-36, Ano87-37, Ano88r, Ano89x]. Granted [Aus60]. Grants [Sci55a]. Graphically [Lev80]. Graphics [Dow82f]. Grappling [And94]. Grasp [Hig76]. Grass [Lak78, Moo93e]. Grassroots [Mac92, Mac89]. gratitude [Ber83, Fer54a]. Grave [Hud70, Ano96m, Kug69]. Gravest [Smi58c]. graveyard [Mil87, Ste89b]. Gravity [Old81]. Gravy [Sie04a]. Gray [Hob85, Kat54d, Sti83, All14b, Roc83a]. Graying [Ste91a]. Grayley [Dow82d]. graze [Ano96m]. Great [Ano49s, Ano49t, Ano53m, Bri86, Dav07, Fel79f, Haq05, Kre90, Kri75, Mar59a, Mar59b, Nee79, Pri53, Rab54d, Rea80, Res76, Sch68f, Sha54b, Sle56, Slo68, Uhl95a, Ano96l, Ano97i, Dro19, Mec18b, Pal77a, Ric79, Rot97k, Sch13a, Squ21b, Hin53, Jac73c, Lar62a, Pat90a]. Greater [Rep86, Ano15c, Bar77b, LaF90b, Lin12, Nat88]. Greatest [Gre64a]. greatly [Alb98b, Gol96a]. greatness [Kes89]. Greece [Ark87d]. greedier [Zag95e]. Greek [Ano68d, Ano68j, Gom75a]. Greeks [VT73]. Green [Ano91-37, Ano92w, Ano96j, Bal79c, Bor71, Dow83d, Fly00c, Rei78, Swe86, Ano90x, Kel90, Lan90h, Sch90, Sob21, Ste89a, Ano87q, Arm68, For09, Kee56, Rob91a, Sar72, Sim65, Ure56]. Green-eyeshade [Ano91-37]. Greenback [Ano96l]. Greenbaum [Sac72b]. Greenberg [Bri86, Nei82a, Sno68b]. Greene [Gri83a]. Greenhouse [Lan90j, Smi90a, Win81a, Kel90, NA92f]. Greening [Rab71b, Han91, Isa90x]. Greenland [Moo25]. Greenpeace [Ano90r, Dan91, Mas90, Rot94e]. Greenwood [Tsi72a]. Greg [Moo95e, Dro21a, Sto19c]. Gregg [Woo92]. Gregory [Dow84d, Sig85b]. Grenville [Sha59, Shi50]. Gresham [Ano64j]. Greta [Dro23d]. Grid [HVK^{+}06, Dro21a]. Grid-Locked [HVK^{+}06]. Griffin [Dow80g]. Griffiths [Roc81, You66c]. Grigori [Mar91a]. Grinberg [Hol86a]. grips [Yor75b]. GRIT [Osg80]. Grobstein [Tho80g]. Grodzins [Rab64c]. Gromyko [Cra88, Ano47f, Bro79c, Gro47a, Gro63, Rab47j]. Gross [Kra94]. Grosser [Mul72]. Grossman [Dow84b]. Groueff [Smi68]. Ground [Dow84b, Fai80, Fed70, Gol69c, Kre05, Pol93, Sea63a, Ste87, Ano88k, Ano14d, Ano18, Cle11, Gre18, Man21, NA88b, Ros98, Tsi87, Ano78n, Ano82o, Dow81d, Dow82h, Kal82a, Kal83c, Dow80j, Mor61a, Dow82d]. Ground-based [Ste87, Ano18, Gre18]. ground-force [Ano88k]. Grounds [Ree07, Str56]. groundwork [Twi23]. Group [Ano52r, Ano57k, Ano58c, Ano60a, Ano60c, Ano60d, Bet75b, Kel79, NA91e, Orr56a, Rab59l, Scl79, Shi47a, Tod79, Ano77l, Ano82c, Ano57j, Ano97r, Ber84a, Bur74, Car72, For83, Fre84b, Sai72, Sch80e, Siv80]. Groups [Ano50-36, Mow83]. grouse [Men91c]. Groves [Gol88b, Ano45i, Ano46-31, Ber03, Gol92b, Gol92c, Gol95b, Gro49, Nor02b, Nor03]. grow [Dro22b, Woo98d]. Growing [AK04, Ano86i, BB10, Car82a, FU63, Fur81, Kat57b, Kla90, Met83, Mus82, Pra06b, Bre13]. Grows [Foz61e, Bar77b]. Growth [Ber72b, Col74, Har72a, Lew56, Mye87a, Myr72, Nor74, Opp59a, Pla70, Sax51, Str67, VT73, Wu99, vD78, vA73, Abb12, Flo81, Hun04, MM86, Rot77, vHW76a, Dow80g, Ber72b]. Gruber [Ded62a]. Grudging [Loc89a].

Grund [Thi46]. grunt [Ano97n]. Guaranteed [Rot61]. Guard [Ano78g, Aur71b]. guarded [BP95]. Guardians [Tra90]. Guarding [BES02, Sie80]. Guards [Ark01b]. Guatemala [Woo97c]. Gubser [McC54a]. Guenter [Kre85a]. Guernica [Rho13]. guerre [Hal93a]. Guerrillas [Hym71]. Guess [Ano96j]. Guest [Arn97a, Ber95, Bla97, Bro62b, Bun98, Cra96, Cum94, Dre87, Fel74a, Gey94, Gla60, Her95a, Ing55a, Ing74, Joh97c, Kre97, Lea97, Mac96, Mak95b, Men95, Mey94, Mur51, Myr80, Obe74, Osb54, Pap98, Pei53, Pik96, Ple96, PY98, Pri74, SL87, Sha96, She97, Sim95, Smi95a, Smi98, Sup61, Tay96, Wei94c, Wil90, Zam98]. guests [GP92c]. Guidance [Ano78g, Dow82f, Pit91]. Guide [Ano85d, Ano89g, Dow80e, Dow83d, Gar83a, Gra84b, Gun66, HB69, Hir59, Kli81, Mes84a, Pre96, Smi83, Van03, Wes74b, Woo01b, Ano84m, Ark95e, Bre85, Kea24, MK02a, Oze92, Rin84c, Ano87q]. Guided [Haf47, Sto90]. Guidelines [Row77, Spi71c, Nov77b]. Guillemin [Adl69]. Guilt [Com54, Gla76d, O’B48, Col93a]. Guilty [Fly01c, Sab76, Ano94p, HWG95, Mak94a]. Guinea [Dow82d, Tho80g, Ack90e, Fin92, Ise01]. gulag [Isa92i]. gulags [Zag95a]. Gulf [Hal93a, McN63c, Ack87c, Ack91b, Alb98d, Ark93b, Arn91, Fot91b, Isa91h, Kle91, Lop91, LC98b, Luc90b, NA90j, NA90u, Wur91a, Ste86c]. Gumption [Doh57]. Gun [Fly98a, Har92a, NA92a, Tor77a, Tra71, Alb81, Uhl95b, FBH81]. Gunbelt [Gol92a]. gunman [Har92a]. Gunnar [Kam71, Sch69a, Myr72]. Gunned [Sch00c]. Guns [Ano78o, Ano92t, Gol92a, Har92b, Nel93a, Str06e, Vid90, Wil99, Ano93w, Ano94t, GL95, Khr97, Zag93b, Zag94c, Bul03a]. Gunther [dF79g]. Gurtov [Uhl98b]. Gus [Ano11c]. gust [Lag80]. Gusterson [CB95, Moo97a]. gut [Day77j]. guys [Zag94c]. GWEN [Ano89j, NA94e, SD88]. Gwinn [Ber78a]. Gwyneth [Ano15c]. gyros [OP98].

H [Abb85, Ack87a, Ano47w, Ano49w, Ano51o, Ano78n, Ano89f, Ano91q, Ber77b, BGR61, Bre87, Bri86, Bri89, Bri80b, Cah94, Cha74, Cle74, Cof66, Div79, Dow88, Dre82b, Eck78, Eng91, Fal61b, Fel81a, Fis83, Fre95, Goo84, Kat52b, Kre90, Kri85, Lof51, Mar78a, Mar89e, Mye87a, Ost72, Pan00, Pos79, Rab60h, Rec89, Sar85, Saw79, Scl79, Smi90a, Smi83, Som84c, Spi66a, Tuc97, Why62, dF79a, dF79d, ABB^{+}50, Ano50m, Ano53u, Ano54i, Ano55r, Ber84c, Bet90, BHM91, Car03, DeW79, Dro24e, FR76, HM90a, HM97, Hot76, Ing54a, Ing54b, Kir55, Kna54, Kno90, Lap56b, Mac54a, Man84b, MSDP80, Mor50b, Mos85, Opp68, Pas55b, Rab55c, Rab56b, Rid50a, Saw54a, Sim55, SB84, SD56, Str54a, Str54b, Ure50]. H-Bomb [Mos85, Bet90, SB84, ABB^{+}50, Ano50m, Ano53u, Ano54i, Ano55r, HM90a, Hot76, Ing54a, Ing54b, Kna54, Lap56b, Mac54a, Mor50b, Opp68, Pas55b, Rid50a, Saw54a, Sim55, SD56, Str54a, Str54b, Ber84c, BHM91, Car03, DeW79, FR76, HM97, Kir55, Kno90, Man84b, MSDP80, Ure50]. H-Bombs [Sar85, Rab55c, Rab56b]. H. [Ano51o, Car68, Ure46c]. H.R. [Ano50w]. Haas [Hur78a, Rei78]. Habit [Uhl97a, Ste77c]. Habitat [Dix65, Kne65, Mid65, Rab61g, Wid61a, Wid61b, Wol65]. habits [Ano07p].

Hachiya [Bru55a]. hacked [Ive16]. hacker [Ano95k, Alv87, Sam06]. hackers [Ano91v]. Hackett [Ler80a]. Hackworth [Rec89]. Hadd [Ano78n]. Haddow [Ing56c]. Hades [NA89s]. Hadley [BGR61]. Hafemeister [Ano91r]. Hafner [Ano86i]. Hafstad [Ano49u, Ano50u]. Hahn [For80, Hah67a, Hah67b]. Hail [Rab61e, Rot93j]. hair [Arn92c, Mar89f]. hair-trigger [Arn92c]. hairs [Ark95c]. Hairy [Mag88b]. Hairy-chested [Mag88b]. Haissinsky [Eme54]. Haiti [Wer93]. Haley [Ano85c]. Half [Ano82l, Aro56, DM82, Dow80f, MW47, BR90, Eva84, Kau90, Moo95q]. Half-Century [Aro56]. Half-SAFE [Ano82l, DM82]. Half-Written [Dow80f]. halfhearted [Wer93]. Halfstad [Ano52k]. Hall [Ano57b, Gol92a, GP92c]. Halliday [Win90]. Halloran [Ada87]. Hallucinations [Shi54c]. halo [Lan90o]. Halperin [Ack87a, BGR61, Cof66]. Halpern [Ano52y]. Halt [And56, Pau88, Man77]. Ham [Ano81f]. Hamad [Mec16f]. Hamburg [Ano55p, Ano87-28, Shi54a]. Hamerow [Ste91a]. Hamilton [Ano57f, Kil78, Pow89b, Pow90a, The89e, Wou81]. Hamish [Kil78]. Hammarskjold [Rab61c]. Hammer [Wei86b]. Hammond [Wol63, Mor62d]. Hampshire [Abb91]. Hancock [Ano46o]. Hand [Moo99a, Ark84f, Cun88, EB78, McC86b, MJD92, Pfe82]. Handbook [Ano90o, Smi83, Ste77b]. handcuff [Moo91h]. Handle [CLR74, Pow89a, Sau17]. Handling [Sel69c, Wil81a]. Hands [Ano89k, Kur79, Lan90c, Mae01, Per90, Rus52b, Dro23i, Fel80b, Fel81d, Kre86d, Man87]. Hands-on [Lan90c]. Hanford [Gra93, Abb04a, Ano03i, CG04a, Gra93, Nel93b, Nil14, SM90, Ste88, Ste89b, Ste90a, Ste91b]. Hang [Gra98]. Hanging [Gad89, Dav76b]. Hanna [Lyt79]. Hannes [Jac73e]. Hanoi [Wes79b]. Hans [Ano94h, Bar58a, Bor52a, Dav79b, Gus85, How65b, Kre85a, Rab73c, Sig85b, Wei48, Ano85u, Ano92-32, Ano14a, McD62, Dow83a]. Hansen [Ano13d, Han90c, Mar97a]. Hanson [Ano46-44, Bal50]. Happen [Kri75, AB18, Dro21a, Bar80m]. Happened [Ano72h, BEGJ72, Ber72d, Cic74, Dee59, Gan72, Gen72, Jac72b, Kri02c, Ano86u, Ano90-54, Bro86i, Fel79m, Fra90a, Gay98, Isa83b, Kam23, Mul95, Tho86a]. Happening [Hal75a, Mec19c]. Happens [Coo75a]. Happiness [Hir72]. Happy [Ark87e, Rot97g, Ano95j]. Harassment [PF75, Ano82c]. Harbor [Car72, Sai72, Mar66a, Par66a, Ure45, Wig66, Ols62a]. Harbutt [Boy87]. Hard [Bet69b, Cha62, Fel77e, Fel57, Har79b, Moo94f, Mur75, Rep03, Sim56, Swe88, Way65, Ano90-34, Ano90-30, Ano95m, Ano97j, Ano07p, Cai88, Eps97, JS97, Kul88, vHJ96, Her83a]. Hard-nosed [Fel77e]. hard-target-kill [Kul88]. harder [Joh94f]. hardest [Abb98]. Hardin [Ano85b, Hul78, Lew72a, Lew73a]. Hardly [Fly06a, Law90]. Hardtack [NA92f]. Hardwar [Lar62a]. Hare [Cal60a]. Harlow [Dra69]. harm [Fre91b, Gor96f, Lam83, SMRVS22, Sie05d]. Harmony [Pop52]. Harnessing [Ano13f, CG12]. Harold [Dow79e, Win81a, Ano47w, Coh81, Kee56, Ure56]. Harper [Wil71, Mec24]. Harriet [Cra83b, Dow84d, Nad79, dF79f]. Harris [Fel82h, Fre95, Kal82h, Sut83]. Harrisburg [Pat79d]. Harrison [Mad79, Mit54, Pla57b, Rab71a, Wil68b, Ano85l, Hol87a]. Harry [Deu83a, Dow81f, Ano56g, Ano74f]. Hart [Mor87c, Pla57b, Gan83c]. Hartung [Moo95c]. Harvard [For83, Fre84b, Gol95a]. Harvest [Bou97, Wal61, All56k]. Harvey [Fel73a, Phe63, Ste76a, Dow82d]. Harwell [McK51, Pfe51, Lon52b, Ski48]. Harwit [Rot97c]. Haskell [Dow79e, Dow83d]. Haste [DeN78, Ial18, Rot95l]. Hastings [Dow81f]. Hatchet [NA88k]. hath [Law75]. Hats [Moo04]. Haunts [Mar61a, Ros89a]. Hauser [Ano62l]. Havana [Ano94s, BT94, Men93a]. Havel [Gol93d]. Haven [Fly02d, Ano81h]. Haven’t [Ing55a, Rot92h]. haves [Rau91]. Having [Isa99a, Plo86]. Havoc [Spi54]. Hawaii [Alb78, Alb81, Jon91a, MJ68]. hawk [Isa88e, Isa96a]. Hawken [Sto17c]. Hawks [Dri93, Isa89k, Isa90h, Isa95a, Isa95b, Gol93c, Wal85a]. Hay [Kri85, Rot93m]. Hayes [Ano78m]. Hayhoe [Ano13e]. Haynes [Rau91]. Hazard [Arn55]. Hazardous [Ano93s, LI70, Har89a]. Hazards [Ano50l, Ano91q, Bri57, Bon53a, CEE85, Com57a, Gla56b, Gro82, Hol74, LeR50b, Lis46a, MD64, Mic74b, Pie61, Pra05d, Rei68b, Ric77, Rot81a]. hazy [Bro84a]. head [You92, Dav65, Hed65, HH65]. headaches [NA91g]. Headed [Ant74, ADK91, Dro25b, Fel77e]. headlines [AH92c, Isa89b]. Heads [Rot92i]. headwinds [Ald17]. Heady [Isa90i]. Heal [Ben69]. healing [Dro23e]. Health [Ano74i, Ano78k, Bor70, Cle07, Fri72, Hia87, Kap76, Lee73, Man04, McC93, New73, Ney73, Pre96, Sci55a, Ter58, Ahe90, Con90b, FWCL20, Gal90b, Ger21, GG16, Hag77, Har89a, Lou90, McC90, OW90a, Per90, RR81a, Rot93w, Sil79b, SS24, Sol79, Ste90a, Ste90b, WvH86, vV23, vHC86, Ano91-32, Hag77, Rad83, Sha90]. Health-Pollution [Ney73]. heard [Alb98c]. Hearing [Rot03c, Joh94b]. Hearings [Ano45e, Ano45l, Ano45o, Ano46-32, Ano46-57, Ano46-92, Ano49t, Ano52h, Ano52l, Ano53a, Ano53b, Ano53c, Ano54v, Fes73, Hol58, Ing60a, Lap59e, Sai72, Sch72a, Sim56, Whi48b, Isa90s, Pat79g, Ano55e]. hears [Sut97]. Heart [Cof65b, Lyn78]. Heartland [Ano89f]. Hearts [Aus99, Eva04, Ray89, Zhe95]. Heat [Bur06, Dro23j, Smi90a, Ano13d, Kor17, Lon52b]. Heated [Dro23m]. heaters [Maz81]. Heathcote [Fin65]. heating [Hir76, Shu76]. heats [SF81a]. Heaven [Pei80, You78c, Ano93t, Ano96j]. Heavens [Hol86b]. Heavy [Ano86j, Ark99b, Tho85b, Cun88]. Hecht [Bro47c]. Hecker [Ano13c, Mec20g]. Heed [Rab69d]. Hegemons [Cum02]. hegemony [Dub85]. Heilbrunn [Cro65b]. Heinonen [Ano11d]. Heinz [Ano56d, Kap61]. Heisenberg [Cas92, Gou47b, Hei54, LRD^{+}91, Pow93, Swe93, Wal90b, Wei93, Swe93]. held [Ano58l, Ano64p]. Helen [Ano78n, Ber79, Dow82c, Dow83b, For80, Cal86a, Jr18]. Helix [Tho80g, Rab68a]. hell [Ack92, Wil89]. Heller [Sag86]. Hellman [Cra88]. Hello [Iva02]. Helmets [Hul98, Ano94v, Thr92]. Helms [Isa88f, Kre97, Whe90]. Helmut [Bar88, Sel63]. Help [AM74, Cum02, Day74b, Dre74, Gol77, Ken61c, Lor74, Rus53, Sal80a, Smi95a, vH00, Ano95n, Dal17, Dw77, Eav17a, Fai97, Gau24, Han92a, Hes78, JH16, MR21, Mil89a, Pie22, SMRVS22]. helped [Ano77l]. Helping [Sch74c]. helps [Ano91-38, Ano91-64, Epp20, Sha10]. Helsinki [JS97, Zim81a]. Henk [Dro25b]. Henry [Abb91, Dev69, Dow83c, Fin90, Ing56c, Mos81b, Mul65, You66a, dF79h, Ano15d, Day77h, Moo99e, Wal46b]. Heppenheimer [Tho80d]. Her [Opp03a, Moo01f]. Herb [Dow81i]. Herbert [And83, Ols62b, Wal81a, Zac76, Ano85v]. herbicide [Pfe82]. Herbicides [Gol70a, Kri85]. Here [All17, Ano54u, Col03, Dow80j, Isa00f, Lan01, Lil49d, ME51, Par78a, Pra96, Rub74, Ano93u, Ano98o, Bru96, Cha88b, Jon97, Kar17, Mec14, Sch80a, Sør80, Dow80j, Dow81i]. Heredity [Bea59, SSW^{+}55, Cro67]. Heretical [Rab51a, Rab72d, Rab72c]. Heritage [Wes55]. Herken [Woo92]. Herman [EE85, Fri62, How65a, Mei68]. Hermann [Bal67]. Hermes [NA90d]. Herneck [For80]. hero [Dro22i, Gor76]. heroes [Var96b]. Heroic [Ale67]. Heros [Rab62e]. herring [Ark87h, SC76]. Hersch [War92a]. Herschbach [Ano87-32]. Hersh [Sid69]. Hershberg [Gol95a]. Herta [For80]. Hertsgaard [Wei89a]. Herzberg [Moo99e]. Hess [Wei85b]. Hessing [She79]. Hetman [PF75]. HEU [HNR12]. Heuvel [Ano90o]. Hewes [Ano50-41]. Hewlett [Gre71, Nie63]. Hey [Rot96f, Ano93t]. heyday [Boy91, Wea85]. Hi [Kel98a]. Hi-tech [Kel98a]. Hiatus [Fit69]. Hibakusha [Sal80a, Tak84]. Hibbs [Alb89b]. hibernating [NA90d]. Hickenlooper [Con48, HL49]. Hidden [Dys75, Han90c, Zag91a, Ack95, Kre86d, Pow81, Rot98f, Joh97a]. Hide [Coh02, Ham83, AH91e, Kre88c, Beu49, Dow83e, Shu49]. hide-and-seek [AH91e]. hideaway [Ano95i]. Hideki [Ten75, Toy82]. Hiders [Kat61a]. Hiding [Tsi72b]. Hierarchical [Pla70]. hierarchy [Ano88k]. High [A.54, Azh93, Cor58, Cut75, Dje68, DSZ74, Fry56a, Isa88c, Kin86, Kre87a, Llu99, Mar70, Mei90, Mil88, Mor61a, Mos90, O’C87, Pau70, Pra92, RS88a, Ric96, Sch75a, SR81a, SF94, Sie85, Wal90d, Wei56b, Wor06, YY74, Ada86c, Ano93r, Ano97j, DM81a, Dro22a, Dro22c, Fel79j, Fie21a, Isa91h, LN86, MFN17, MW91, Ram16, Ray80, Tel47b, Tir85, Win79, Hin60, All55l, Deu83b]. High-Energy [Pau70, Wei56b, Tel47b]. High-flying [Pra92]. High-Level [Mei90, SF94]. high-pressure [Fie21a]. High-tech [RS88a, Wal90d, Dro22a, Dro22c, Isa91h, MW91]. high-temperature [Ram16]. High-Voltage [Sch75a]. Higher [Ast67, For70, Hes84]. highest [Abr83]. highlight [Fri88]. Highlights [Ano50v, Ano97s, Kel10]. Highly [NA94e, Rot93a, Kap90b, SD10b]. Highway [Mos92]. Higinbotham [Ano22c]. hijackers [Sob91a]. Hijacking [Wri99b]. Hill [Gor60, Ano74d, Cas77d, Mar64f]. Hillen [Hul98]. Hilliard [Bla84]. Hillman [A.52]. Hillsborough [Cro65a]. Hilton [Ano94l]. Hilts [Gen84a]. Himself [Rev66, Gen84c]. hindering [Spr25]. Hindu [Sha98a]. Hinge [Lor81]. Hippel [Ano15b, Day75c]. Hippocrates [Ste69b]. hire [Alb93c]. Hiroshima [Bro60a, Bru57, Bru59, Cog82, Gol95a, Hig66, Kre90, She83a, Alp92, Ano90j, Bar77b, Ber75b, Dow80h, Dow82f, Dow82e, Dow82h, Dow83b, Fel85b, Fri95, Fri20, Hos86, JR62, MSW20, Mor52, Nat77, NK86a, OJCP^{+}47, Osg85, Pop81, Per13, Rab46b, Rab47m, Ric12, Sag71, Shr70, Fat46, Swa86, UC62, Wei85a, Mor62b, Bru55a, Cog82, Dow82e, Mes83a, Moo95e]. Hiroshima/Nagasaki [Dow80h, Dow82h, Dow83b]. Hiroshimas [Dow83f]. Hirschfelder [Fel81a]. Hirshorn [Mes64a, Dow84d]. Hiss [Dow81i]. Histopathology [War49]. historian [Hew81, McN82]. Historic [Ano59e, Ano03i]. Historical [Ahn10, Ano96q, Ben56, Chi78, Haw47, Par46b, Pit91, Ros60, Boy86a, Osg83]. histories [Edw12]. Historiographical [Wal90a]. History [Ack88d, Ano49v, Ano51t, Ano58a, Ano91k, Ber93b, Bet90, Bor52a, Bur09, Bur47, Cum01a, Fel70b, Fel83a, Fer69a, GP84, Gen84c, Git54, Gre64a, Han98a, Haw50, Hig63, Hig66, Hol86b, Ise92, Lit54, Lor81, Mar59a, Mau50, Mos91, Nac72, Nie63, NCA85, Oom46a, Pip55, Rab56g, Rab57g, Rab63h, Rau91, Rot97c, Sak69, She83a, She89, Smi88, Bul02d, Sto59a, Swe93, Ure69a, Wei48, Whe84b, Win84c, Win84a, Win94, Win98, vL92, Abr83, Ack95, All16, Ano90i, Ano91-60, Ano91j, Ano97k, Ano03i, Ano14d, Bez84, BS97, Cal86b, Car03, Dav86a, Gow79, GH83, GJ84, Har96, HK02, Hod81b, Koe77, KN17a, La 91, Lan89d, Les93b, Mar82b, Mec20g, Nye13, Oli89, PN09, Pow81, Pow93, Ram16, Ric84, Rot93p, She87, Sim11, Sme22, Smi84b, Sta00, Thi46, Tsi84b, Wal21a]. history [Wei83b, WB85b, Wu23, YG77a, Cum98, Gar97a, Haw56, Ost72, Rot97b, Sch81a]. Hit [Kyo85, Tsi75a, Isa95b, Man21]. Hitch [Bou61, Zha15b]. Hitler [Gou78b, Ano95i, Kin69]. hits [SC97, Van86]. Hoax [Gre55a, Kee56, Ure56]. Hoban [Mes82]. Hobbled [Riv62]. hobby [Ano92x]. Hoc [Tod79]. hocracy [Lov74b]. Hoenig [Cal84a]. Hoffman [For80]. Hogan [Ing70]. Hogar [Ano91u]. Hoggart [dF79a]. Holbrooke [Sha98b]. Holcombe [Sto59b]. Hold [Gab01, Sig01, BHK^{+}21, CS16, Isa92c, Kre83b]. Holder [Ano98j]. Holding [BW18, Ric03a, Ols16]. Holdouts [Bar10a]. Holdren [Ben07, Eck78]. Holds [Mot48, Noy47, Col88, Zam93]. Hole [Pai04b, Dro23e, Men89]. Holes [Utr87, WB87, Ano86z]. Holiday [Ano46s, Sch05a, Woo98b]. Holifield [Lap58, Sim56]. holistic [Dan16]. Hollaender [A.54]. Holland [Smi76]. Holloway [Sta85, Cra84d, Fre84a]. Holly [Lan62]. Hollywood [Ano92u]. Holocaust [Gid73, Kre98]. Holocausts [Bri87, Hos87a, Hos87b]. Holroyd [Ano85c]. Holton [Saw80, Bar66a]. Holy [Rot93k, NA91l]. Home [Ano90-30, Bos06a, Fly00b, Gov93, Gra93, Gre87, GM93, Hau68, Lap59d, Mei77a, Rab68f, Sie05c, Sim54e, Wal69, Whe84b, Ano91-50, Ano94k, Ano95i, Gom75b, Joh89, Joh96a, Lor00a, RL89, Sut89, Woo98c]. Homefront [Sut96a]. Homegrown [Rey04]. Homeland [Kar17, Sch05a, Sie04a]. homeless [Zag94c]. Homer [Aft01, Ano12f]. Homer-Dixon [Ano12f]. Homes [Aue06a]. homicide [Nas80]. Homo [Bor52a]. Homogeneous [Haf50]. honchos [Men93a]. honest [Fel82c, Fra47]. honeybees [Gif13]. Hong [Gal91, Gom75b, Hum83, Ste97]. Honor [Ano63w, dF79g, Pee78]. honorable [Ano94n]. Honored [Ano72f, Van89]. honour [BM99]. Hooked [Lan89b, LP92]. Hope [Amr54, Ano78m, Ano86-35, Ano89r, Ano95j, Ano96m, Bor64, Bre93, Dav77c, Dro24e, Hig66, Kos60, Shr70, Tay47, Tel46b, Wie68, Zei64, Ack87i, Che89, Cla77, Day96, Eps86, Eps91, Mec19g, Moo91c, RK91, Smi65, Smi70b, Wei79, Gom84a, Swe88]. Hopeful [Moo98a]. Hopes [Mue01, TvH83, Zel76a]. Hopi [Yoa78]. Horelick [You66c]. Horgan [Gol96a]. Horizon [Ber63, OR70, NA87j]. Horizons [Gle51, Gom81c]. horizontal [Fel78k]. horn [Ste97]. Horne [Sim87a]. Hornell [Pla57b]. Hornet [Mat98d]. Horowitz [Nei82a, dF79f]. Horror [Bar80e, Mas99a, Rho13, McK23b]. Horrors [Joh97a, Fre95, Pin24]. Horse [Sap75b, Cap77, Com76b, MH89b]. Horseback [Wri63a]. horses [Abb89]. Hose [Dje75]. hospital [Abr83, McC86b]. Hossein [Ano15g]. Host [Llo71]. hostage [Fel78g, Myr80]. Hostages [AH92c, Ram86b, Ste82a, Gol83c]. hostility [Bos88]. Hot [Ano94n, Ano07h, CO74, Gro06a, Isa02d, KKL15, NA90c, She83a, Ano91-53, Ano92q, Ano95i, Ano95k, Ano97k, Ano14b, Eps82a, Sch92a, Ste91b, Zim89c]. Hot-Air [She83a]. Hotel [Hal93a]. Hotline [Bla84]. hotspots [Zac82]. Hott [Dow80e]. Hough [Val86]. hour [Shu78]. House [BBG^{+}85, Cla66, Gaf86, Hor77, Rab68g, Sue68, Zei54, Bli76, Sin83, Ano92t, Che52, Chu54b, Chu54c, Dow83c, Fla78, Isa86c, Isa88c, Isa90j, Isa93d, Jac74a, Lal66e, Moo98h, Shi46a, Shi48b, Whi66a, Whi66b, LLR^{+}47]. Houser [Bar90b]. houses [Shu76]. Housing [Nan73]. Houston [Wyl56]. Hovers [Hal75b]. how-to [PC92a]. Howard [Fra91, Gre82b, Pre96, dF79g]. However [SCHD20]. Hoyle [Pos79]. HP [Ano94u]. HP-3 [Ano94u]. HP-6 [Ano94u]. HP-9 [Ano94u]. Hu [Pre96]. Hua [Moo99b]. HUAC [Boy91]. Hubertus [Wri63a]. Hubris [Rot03a]. Huddle [vH77a]. Huddling [Men93a]. Hudson [Sin79a]. Hugh [Fal61b, Moo97a, Rep85a, CB95, Wol50]. Hughes [All59l, Cha58b]. Huisken [Siv76]. Human [Abr87, Ack90e, Adr54, Aja78, Alf81, Ano72h, Ano78n, Ano78p, Arn54b, Bal79b, Bix49, Bok49, Bra58, Bri49, Bro75, Car55, CKA10, Dow80e, Dub70, Dum80, Dys69b, Dys69c, Ehr82b, Eis66, Eva82, For80, Fre72b, Fri72, Fri79, Gid73, Gol78c, Got69, Haw58a, Hel77a, ILA70, Jor67, Kan81, KK72, Kri85, Led66a, Mad79, Mar49b, Mar64j, McD74b, Mel75, Mel78, Mor59b, New73, Opp59b, Orn67, Pla57a, Pla73, Pri81b, Rab68d, Rad83, Rob77, Sab72, Sag73, Sag86, Saw81b, SSW^{+}55, Say55, Sin56, SD55, Soy89, Sta49, Sti83, Tho80g, Uhl98a, Von83, Wei66c, WG69, Woo82, Zim78, Zim79, vL86, Ano87-38, Ber82, Blo85, Bos98, Cro17d, Dav77a, Gol82a, Gra86, Hai86, HFMS79, Lop76, Man87, Mei77a, PO16, SMĆ08, Sco92, Ste81b, Ste82b, Was82, Yor75b, EV84]. Human [Dro22a, Nor64, Fri84, Mac83, Smi84a, Hul76, Wyl56]. Human-Climate-Wildlife [CKA10]. Humane [CF74, Arn80]. Humani [Pop52]. Humanism [Cha73a]. Humanist [Fre67, Mei73, Spe66]. Humanistic [Bro68]. Humanists [Har79b]. Humanitarian [Bac03, Lap56b, Bor14, GG11b, Mul95]. Humanities [For80, Gil84, Com81]. Humanity [Mul59, Rot96a, All15a, All18, Bar80h, CW23, Lif13, Tod81, Par86]. humans [Ban22, CW23, Nuc20]. humming [AZ89a]. Humor [Rot01c]. hump [Kam20]. Humphrey [Eli85, HM72, Sim59f, Spi57b]. Hunch [Dud75a]. Hundred [Pla57b, Ark96e]. Hung [Mar94]. Hungarian [Gya90]. Hungary [Ber89a, Van90b]. Hunger [Ano86h, Bor70, McG67, Mul53, Orr56b, Bra86a, Bra86b, LB78a, Oko82, Bra87]. Hungering [Sch67b]. Hungry [Str07c, Fel76e, Sch77b]. Hunt [Ano99-31, Dow88, Kor92]. Hunter [Nel84]. Hunting [Ano07h, Ure49a]. Huntington [Mor62d]. Huntzinger [She79]. Hurricane [Ada73]. hurry [Ano90-33]. hurt [Dal17]. hurts [Tho84b]. Hurvic [Pol68]. Husbands [Rep86]. Hussein [Ark00e]. hut [Ano93x]. Hutchings [Bul09g]. Huts [Rot89a]. Huxley [Doh58, Doo60, Lut73, Mor59a, PMRP50]. Hwang [Uhl98b]. Hybrid [Chi17, For09, All17, Dal17, Mec17c]. hybrid-war [Dal17]. Hydrogen [Arn50, Bac50, BBSS50, Bet50, Bet63, Dea54a, Han95a, Hay55, Rab52d, Rab54d, Rho95, Rot55, Thi55, Ano53o, Rob78b, Air55, Col54]. hydronuclear [AS94]. Hyman [Ano49w, Ano52-28]. Hymes [Hym69b]. Hynek [Val73]. hype [AH91a, AH92c, DeW88, Gou23, LN86, Mec23d]. hypersonic [Gub15, Nag15, NA87e, Oel20, Zha15c]. Hyperwar [Arn92d]. Hyping [AH91b, Oel20]. Hypocrisy [Pow04]. hypotheses [Gar97a]. Hypothesis [BW77, Wei77b]. Hysteria [Spi71a].

I. [Ano45j, Ano88m]. I.C.C [Ano45c]. I.C.C. [Ano46-45]. Ia [Hol86a]. IAEA [Adv62, Ano86-37, Ano88v, Bol88c, BOW10, Ekl77, Ekl81, FM82, Fis82, Flo88, Hig89, Jon93b, KSG15, Mog81, Roc18, Rod61, Rub65, Van90a]. Ian [Dow80f, Say84]. IARPA [Eav17b]. ICBM [Bot69, GW92, Jac74b, Kea60, Mor84b, NA87i]. ICBMs [Hob85, Ark84c, Tal18]. Ice [Fle70, LH70, Spe74, WZSA73, Woo73a, ZSA73d, Bra17, Dro24g, Jac80a, Mer24, Moo25, Sca11, Sie24]. Icebergs [Lon74, WC73]. Iceland [Ano86-35, Ark00a, Coo78]. Iceworm [NA89c]. Icon [Hel07]. iconic [Hus24]. Idaho [Pau88, Sie05c, Sta00]. Idea [Ano86h, Ark99c, Con67, Cum94, Jon63b, Rob08, Sko67, dF79a, Ano96o, Ano96z, Koe22, LP97, Rot98c, NA90d]. Idealist [Cor53]. Ideals [Lan90n, Sal76]. Ideas [Dil66, Ein54, Fre63, Hut53b, Lan92a, Rot03b, Bel20a, Ger82, BB63]. Identification [Bro60c]. Identity [Gra61, Mil69c, Mot89, Ber81, Lüb90, Mar89b]. Ideological [Had77, Pow90c]. Ideology [Sch70b, Vuc69, O’B82, Fra91, Par88b, Dow88]. Idle [Com74]. IDRC [Spu72]. iee [Vog90a]. If [AS02, Ano92r, Ano98n, Bou74, Bus47, CW23, Com46, Dav79c, Dys75, Fel80d, GB11, Gla75a, Har84b, Lor02, Mar82c, May90, NA91m, Pod05, Rab47d, Rot83, Sha92b, ZRG83, Ano97k, Che52, Cot22, Dro19, For02, Gla85, KK18a, Pie22, Roc18, Sch16, Set16, Sim16, STRS80, Tho84b, vH79, Fra80, Dow83b]. ifs [Obe97b]. Ignoble [Ano96k]. Ignorance [Lee60, Pan71, Ves05, Wil81c]. Ignored [Ada47, Ack88f, LP91]. Igor [Rab67f, Hol86a]. IGY [Cha60, Sul58a]. II [Ada52, And65, Ano88d, For80, Gre67b, How65b, Pop52, Aft01, AH91a, Ano47m, Ano49d, Ano49c, Ano60a, Ano60c, Ano60d, Ano60b, Ano69n, Ano71c, Ano73f, Ano79g, Ano80i, Ano89c, Ano97m, Arb93, Ark83d, Ark84f, Art71, Bec79, BS47, Bor65b, Bro79e, Bro85h, DB52b, Dan54b, Dow83d, Dre77, Ein79a, Ell92, Fel72b, Fel73d, Fel80k, For51b, Fry65, Gal56, Hal66b, Hum72, Ing56a, Isa78, Isa01a, Jac72b, Kal61, Kli52, Kno60, Kor90, LR79, Leg56, Loc92b, Lon52c, Mil81, Mor51, NA91e, Pai80a, Rab56l, Rab72d, Ram72, Ros89b, Sha78b, Sha00, Ski53, Tri71, Tro79, Wid61a, Joh97a, Kre89c, Kre98, dF79a]. III [Ano78n, Gol88a, Gre67b, How65b, Ken79, Ano60c, Ano69o, Ano74h, Ark98d, Aug56b, Bor65c, Bro86f, Buc60, DB52c, Dru71, Dub56b, Fel72a, Fre72b, Gen72, Gro56a, Had56, Han96a, Isa01a, Kob70, Lap55e, Lap73b, Lar65, Lon52c, Mot51, Pri89, Rab51j, Rab56g, Rya89, Wid61b, LaF90a]. IIl [Fel78m]. Ike [Sha91a]. il-Sung [HC12]. I’ll [Ano93z]. illegal [BV20, MSW20, Mey85]. Illesh [McC88a]. Illich [Hag77]. illicit [BWL21, SL13]. Illinois [Gus70, Jac73b, Jac74d]. Illogic [Har85b]. Illumination [AM74]. Illusion [Cav04, Nie49, Pie62, Eps87b, Gei81, Rat92, Reh14, Ble87]. Illusions [Ikl76, Mus80, Per71, Som05, Cul79]. Illustrations [Chi78]. Ilyas [Gau24]. Image [Myr72, Tho80g, Fal89]. imagery [Moo95e, PPE22]. Images [For80, Smi88]. Imaginary [Moo95b, Per91a, Ust93d, Wat77]. Imagination [All06, Den53, Bri89]. imagined [Wal21a]. Imai [Mos81b]. imbalances [Mat82a]. imitate [She85b]. Immaculate [Ano94l]. Immediate [Ano49-42, Lon64, ASB^{+}84]. Immigration [Ano55b]. immobilize [Dah77]. Immoral [Ark00f, Per85]. Immorality [MS85]. Impact [Ano69p, Ano78o, Ano78p, Ano84k, Bat69, Bec55, Ben71b, Cav53, Gee66, Gla75b, Gre62, Lal64f, Lon69a, Mei69, Mor55, Net70, Rab59f, Sch58a, Sch75a, Sco71, Wig51, YY74, GG16, Gil80, PO16, Sas17, Sch77a, Wes71, Sak69, Tsi72a]. impacts [AB18, Coh76d, MM11, PE24, RT12]. Imparsible [Rob91b]. Impasse [Ano96i, Had77, How65b, Sab63, BW94, DPT17, Dro23g, TvH17]. impedes [DM85]. Impending [Rab46g, STC74]. impenetrable [WB87]. Imperative [Ehr71, Gus85, Kil65, Bet85, Jon84, Sco08b, War78]. Imperatives [Fri72, Hai62, Law88, Moo93h]. Imperfect [Fra74, Pat80c]. Imperialism [Whe84b, Ano93v, Von83, Mac83, EV84, Fri84, Smi84a]. imperil [Isa87e]. impermanent [McK25a]. Impinges [Rap65a]. Implanting [McD74b]. Implement [Mar48e, Mec20a]. Implications [Cow65, Dan54b, Dow80d, FT76, Fla65, Ful69, Mat50, Nor74, Rus56b, Shi46b, Spi71b, Sul58b, TT69, Tuc10b, Bis14, BBL13, Cha11, Cle82, Coo11, DR14, HSJH18, Jor11, Kos10, Lef77a, Mor14, Squ13, Tak17, Woo12, Got50, Got51]. Importance [Gom75b, Kay49, Kis66, Sym50, Fra76a]. Important [All50, Ano47l, Bro47e, Den47, DuB63, Fel62c, Due21, Hol88, WW80]. importers [Har87c]. imports [Tyl89]. Impossibility [Rav65]. Impossible [Mor56, Sok01, Ure48a]. Imposter [Ano78n]. Impressions [Ano66k, Coh75a, Lon52c, Agn82b]. improved [Lan90f]. Improvement [Jou52, Spi49, Wei52a]. Improvements [Ano60a]. improves [PPE22]. Improving [Bor14, Fra74, Hak62, MD66, Swa66]. Improvised [Lev06]. Imprudent [Ken80]. Impulse [Lor57]. Impure [HR51, Wyl51, deF51]. in-comprehensive [Joh96b]. inability [Ano94l]. Inaccurate [Ano48s]. Inaction [Mar06b]. Inadvertent [Bla84]. Inc. [Ano55v]. incapsulation [Ano94l]. Incentive [Fri74a]. Incentives [Fre78, Kir63, Elb81, WH16]. inch [NA89f]. Inching [Wur96c]. Incident [Moo92l, Phi60, FK89]. Incidental [LS58]. incidents [BBP14]. Incineration [Cas02, Smi95b]. Included [MS06]. includes [Rot18]. incompetence [Les93d]. Incomplete [Jor87, Ste60b]. Inconsistency [Hur77a]. Inconvenient [Moo07]. Increase [Foz62e, AH22]. Increased [WH72, Dje76, DS12, Eps95]. Increasing [Ano84j, Jac56, Pod16, GM90, Sto18a]. incredible [Squ17a]. Increment [Ure48d]. incrementalism [Cor96]. Indefensible [Isa05c, Keu83]. Indefinable [Bos06b]. Indefinite [BT95, Eps95, Toy95, Ach12]. Indentured [Nad72]. Independence [Fra91, Sut01, Fre76b, Kas15, Rot78a, You76a]. Independent [Bar73a, Dro22g, Dro22h]. Index [Ano49j, Ano50g, Ano51d, Ano52e, Ano53j, Ano54f, Ano55d, Ano56e, Ano57e, Ano58b, Ano59b, Ano60g, Ano61e, Ano62m, Ano64h, Ano65d, Ano66a, Ano67c, Ano68b, Ano69b, Ano71a, Ano72d, Ano73a, Ano74a, Ano75b, Ano76a, Ano77k, Ano78h, Ano79b, Ano80e, Ano81d, Ano82b, Ano83e, Ano84c, Ano85e, Ano86m, Ano87t, Ano88f, Ano89h, Ano90q, Ano91t, Ano92k, Ano93o, Ano94i, Ano95h, Ano96h, Ano97h, Ano98i, Ano99s, Ano00h, Ano01j, Ano02h, Ano03j, Ano04h, Ano05i, Ano06i, Ano53e]. India [Ano64p, Gan83b, Pal64b, Ahm80, Ahm99, AZ89b, AH92d, Alb98c, Ano70g, Ano02i, Arn96b, BV97, Fot89, Gan72, Gan83e, Gho15a, Gom74, Hoe09, Kap72, Kap74, Kar97, Mak97, Mil89a, MF68, Muk93, Nan71, NA98a, NA98b, NAKH02a, NK05c, NK07a, Pal64b, Per02, Pou78, Que70a, Raj10, Ram13b, Rao87, Red82, Ret99b, RTB^{+}19, RR68, RL78b, San14, Sar72, Sas17, SJ82, Sha69, Sha91c, Sin65, Sto68b, Str71, Sub68, Swe78b, Tar23, Voh70, Wad74, Wil82c, Wit01b, Gan83c, M.84]. Indian [And68, Ano07i, Bar80f, Bey80a, Cha16, Dha16, Kar14, Kaz14, KN12a, KN15b, KN17b, KK18b, KK20b, KK22a, KKJK24a, LG15, Mit58, Nag15, Nar67, NK08c, NK10c, RR15, Sha14]. Indians [Ano93z]. indicate [Ano93w]. Indicators [dF79f]. Indictment [MK79]. Indira [Gan72, GD85]. Indiscretion [Uhl03a]. indispensable [Nor02b, Nor03]. Individual [Bri58, Cle48b, Daw68, Kel64, Sci55b, Sch68c, Wol68, Wyl56]. Individualism [Pol49]. Individuals [Gra84b, McK23a]. Indo [CS17, Hym71, Kum04, Per95]. Indo-China [Hym71]. Indo-Pak [Per95]. Indochina [Sta71]. indomitable [Bau84]. Indonesia [Sut01]. Indonesian [Pag68]. Indoor [HNB81]. Induced [Wil92a]. Industrial [Abr73, Ano46-58, Ano50a, Ano51p, Ano60f, Aug48c, Axe51g, CHP71, Cae73, Eis67, Gol92a, Haf54, Joh60b, Lap51a, LI70, Lil49a, Lon69a, Mel71, Mil50, Pla53, Rod70, She61, Spo50, Tho57, Tsi72b, Vof72, Wol71a, Wol71b, Bre13, Han87b]. Industrialization [Mah59b, Gro15]. Industrialized [Ano91q]. Industries [MTK46, Sch49, Ald76]. Industriosphere [Hal75a]. Industry [All55a, Ano49g, Ano51l, Ano52t, Ano54j, Ano64l, AM74, App46, Bec55, Cic74, Com74, Com75a, Cre70, Foz63b, Gle51, Hil65a, KK02, Lil47a, Lil49b, MvE13, Nel74b, Put52, Sch08b, STFK09, Spo51, SM89b, Wei53b, Abb82, Alb86, Bar81c, Bar81h, Bra13b, E.77, Her78b, Lat84, Mar76, MA83, SFT11, SF12, SF14, Ano83l, M.85, Nas82, Ste95, Bei84, Sin78]. Ineffective [Ano51q].

Inefficiency [Tay76]. Inertia [Cic74]. Inertial [Don82, GG11a, Win81b]. inescapable [NB77]. Inevitability [Pal77b, Per11]. Inexorable [Eps75c, Fel84d]. INF [Ano89g, Som91, Ano87-44, Flo88, Flo89a, Gar84c, Kre87a, Kre87c, Kre87d, Lin84b, NA88c, NA89b, NA89m, NA90d, Sig87b, War88]. Infant [Ano70b, NW70, Sag69, Ste69e, Tam69]. Infants [Ste69f]. Infectious [Cro18, Lwo60]. Infeld [Ber79]. infighting [Isa87d]. Infinite [Led71b, Rom89]. Inflation [Thu79, Sap80a]. Influence [Ada47, Fis83, Shi57, Gre78, Kim22, Bul67, Ash81]. influences [Sim84c]. influential [Har10, Sch75e]. info [Ano92m]. inform [IS15, Nye13]. Informal [Bac77b, Lev89]. Information [Alp62, Ano48e, Ano78o, Ano87n, Ano89g, Ano95u, Bec55, Dub62a, Gla62b, Gre68a, Hak62, Hug56, Lal64d, Liv56, MT55, Nor54, Rab52h, Rab53d, Ray81, Rig65, Sal53, Sch63d, Sem53, Sim54d, Ano96y, Hal93a, Hof86, Lin19, MD24, Sto18c, VU22, Ano45h, Ano46-94, Lev51b, Spi50]. Informational [Ric74]. Infrastructure [AF85]. Inglesby [Cro18]. Inglis [Mer79, Rub73, Sto17a, WI71]. Inherently [Say93]. inheritance [Hos86]. inhumane [Gol83a]. Inhumanity [Hor76, Doy19]. Initiated [Ric74]. Initiative [Bot61, Fre84a, Gor51a, Kre87f, Ano76d, Ano79j, Ano85i, CO85, Daa87, Fah90, Isa91d, Mit97, PG79, TT83, Tru84, Uni78, Wil82a, Wol93, GD85, Sah20]. Initiatives [Lev63, Was63, Bar80j, Har84c, Hui87, Lon84, Sha78a, Sha78b]. Injected [Ano99t]. injurious [Hag77]. injury [Bar77b]. Injustice [Asi57]. Ink [Arb93]. Inman [Pai82a]. Inner [Ano78n, Che70, Ede89, OR70, Sha95a]. inning [Men96b]. Innocence [Ama77, Com54, Wil75c]. Innocent [Fly01c]. innocents [aI98]. Innovation [Ano64k, Bei84, Lal67g, Mic64, Sei71a, Dha16, Huq16, Kam14, PO16, SW16, Sch79b]. innovations [Ano86t, Bro85a]. Innovative [Myn11, KK18a, Mec20e]. Input [De 74]. Inquiry [Ano49s, Ano49t, Dys07, Fay75, Sch58b, Smi74c, Woh78, Saw80]. Inquisition [Cha73a]. Inquisitor [Whi53]. Ins [Fla65, Joh66, Isa91p, NA91l, Wal82]. Insect [Wil73a, Wri65]. insects [Odh81]. insecure [Ano94m]. Insecurity [Rot02b, Moo92h, She14]. insensitivity [Mot85]. Insider [Lon85c, Kla19, MR90]. insidious [Mcc86c]. Insight [Oli58, Ano96q]. insights [Tsi78c]. Inspecting [NA90n]. Inspection [Gor67b, Hen59, Lal65f, Mel58, O’S63, Ore61a, Soh62, DeV88, GK87, Nat85b, Cav58]. Inspections [dRMT^{+}47, Fis89, Gol91c]. inspector [Ric81]. inspectors [Smi91]. Inspiration [Wig62a, Rho13]. inspired [Ben79b]. instability [Abr87, Hil12, Per95]. Instead [Fie21b, Gei16a, Pie22]. Institute [Ano46-69, Ano50-36, Ano53n, Bak60, HKL^{+}49, Leo65, Coo85, DW85, Bar77c, Bar78b, Bar79a, Bar80b, Bar81a, New67, Tod78a, Wil68c, Hol86a]. Institution [O’K84, Dun84]. Institution-wdie [O’K84]. Institutional [DeW60, KKHK79, Sch67a]. Institutions [TV71, Cox86, Luc82, Ven79]. Instrument [Fel73a, Cra83b, Roc81]. Instrumentation [Ano60a]. Instruments [Ano49g, dC78]. insufficient [Sch88a]. insults [Kap92].

Insurance [Ano46-67, Bro76a, Nit92, You76b]. Insure [Ure48c, Ack92]. Insurer [Oel14]. insuring [Pug88]. Integral [Rab59g]. Integrated [vdB65]. Integrating [KH13]. Integrity [Mei53]. Intellectual [Ćos65, End00, Lev64, Rab68c, Rav65, Tay58, Art85, Hol77]. Intellectuals [Rei74, Sch68g, Shi67, Ein48a, Smi60]. Intelligence [Bam85, Bra58, Cra83b, Eis99, Jen70, Lew70f, Lew70g, Mos91, Pra02, Pra04b, Pra04d, Ric89, Wri48, Ark87f, Ded78, Fla80a, Gei17, HD23, Hib95c, Mec19b, Pra89b, Pra97, Ric86b, Sep23, Sha18, Tor19, Mec23d, Asp79]. Intelligent [Ano87s, Bel93, Eis99, Vog13]. intensity [Bar90a]. Intentions [Ade72, BV97, Bro85f, Lov86]. Inter [Cou77, Dun84, Fra61a]. Inter-American [Fra61a]. Inter-institution [Dun84]. Inter-service [Cou77]. Interaction [Cra83b, Fed72, Dje76]. Interactions [Hal75a, Tod78b]. interbreeding [Pat80a]. interceptor [Rob18]. Intercollegiate [Ano46-46]. Intercontinental [NK09c, Tsi75a]. Interdepartmental [Ano50u]. Interdependence [Top63, VT73, Hea83, Sny85, Hum66, Cop65]. Interdisciplinary [Fra83, Rin84a, Got20, Ano89b]. Interest [Ano62-28, Ano88e, Dam61, Foz61e, Ham93, Sag73, She64c, Was64, Woo82, Ano15b, Flo76b, Her60, Les93b, Mec21g, NR79, NK15, Sac72b]. Interests [Hol07, Lac85, Bul08a, You66c, Man15, NH08b]. Interference [Dan54a]. Interim [Ano48a, Ano49x, Ano57a, Cav50, RL14]. Intermediate [Smi77a]. Internal [Ano66m, Tow66, War49, Cro67]. International [Ada52, Ahm04, Alb87a, ASB^{+}84, Ano46r, Ano46-27, Ano46-57, Ano46-72, Ano47d, Ano47i, Ano47k, Ano48z, Ano48-31, Ano55l, Ano62s, Ano63p, Ano64q, Ano67h, Ano75c, Ano83d, Ano89g, Arn47, Bak60, Bar77c, Bar78b, Bar79a, Bar80b, Bar81a, Bar46a, BM02, Bla84, Bol63, Boo46, Bor73, Bra47, Cou48, Cav50, Cha60, Chi69, Chi59c, Cle87a, Coc59, Con46b, Cow65, Cra83b, Cra64, Cre63, DS47b, Dee57, Dom82, DSZ74, Ein48b, Ekl70, Elm66, Eme60, Fal87, Fal89, Fed70, FCC^{+}46, Fox48, Foz61c, FK89, Fru61, Fru63, Gal65, Gil55, Gla60, Gol95d, GG11b, Gro46, Gué66, Hai62, Ham61b, Hin18, Hod55, Hoe88, Hol85b, Hui87, Hun87, Hur78a, Ima70, Jas81, Jor63, Kas64, Ken63a, KFM50, Kra46, Lal66e, Liv68, LK10, Lov63, Mal50, Mal85, Mar70, McK74, McM48b]. International [McM67, McN62, McN63c, Mec16d, Mot47, Mye87a, Nac65, Opp46, Opp48a, Opp63b, Opp63c, Osb49b, Pug71, Pol66, Pow59, Rab46e, Rab51c, Rab54e, Rab56e, Rab63e, Rab64f, Rab65h, Rab65d, Rip68, dRVT^{+}47, Ros63, Rot92v, Spe58, Sab63, Sal68, Sch60b, Sch62a, Sch63e, Sea70a, Sea63b, Sea67, Sei73, Sha87d, Shi48b, Shi48c, Sho46, Sim84c, Smi69b, Smi70c, Smy60, Ste76a, Ste73b, Sto72, Str54a, Str84, Sub87, Szi50a, TT69, Tod64a, TV71, Tod79, Top58b, Väy87, Wad63, Wei58, Wei69a, Wen53, Wer73, Whi64, Wil67b, Wri67, You67b, You82, ZSA73d, vdH61, Ano89w, BBP14, BB20, CG12, DPT17, Eps82b, Fai81, Fel77l, Fis86b, Fra75, Hes78, Hud77, Kap81a, Kat80b, Kla79, Kow77, Kri14, McN77, Mec22a, Moo14, Nat85b, Rab50e, Rap11, Sub77, VU22]. International [Ano84k, Bar81i, Bec58, Col59, Dra73, EBC83, Fed59c, Fin65, Jos61, Lil47c, Opp47b, RH65, Rab66a, Ran59, Sac46, Sim57a, Tro58, Wei56b, Wil73a, Zlo47, Ano57c, Ber73c, Ehr82a, Pom67, Rei78, Win84b]. Internationalism [LT70]. Internet [Ark97b, CG12, HV18, Mec21a, Mow95b, NA97c, Zit17]. Interoperability [Moo01a]. Interpretation [Shi48a, Gou78c, Tsi78b]. interpretations [Ano86r, Ric86a]. Interrelation [Soh64]. Interrogates [Ano47u]. intersecting [PFW18]. intersessional [McL11]. intervenors [AL79b]. Intervention [Car72, Gen71, Har70, Bar90a, Kat90b, Mul95, Str81, You91, Gol69b]. Interview [Ano52k, Ano53v, Ano09, Ano13c, Ano14d, Ben07, Dro19, Dro21a, Dro21b, Jr22c, Dro22a, Jr22a, Dro22f, Dro23c, Dro23d, Dro23e, Dro23j, Dro23m, Dro24c, Dro24b, Dro24e, Edi60, Gau24, Jac74c, Krz20, Lap54a, Lap58, McD62, Mec21c, Mec21d, Mec22b, Mec22a, Mec23b, Mec23c, RS52, RDS^{+}52, Rez07a, Rez07b, Rez07c, Sie07b, Bul07b, Bul07d, Bul08l, Bul08m, Bul08n, Bul08o, Bul09b, Bul09c, Bul09f, Bul10, Sto24, Bro20, Dro21c, Dro22h, Fie19, Len21, Mec23a, Mec24, WS21, Dro22g, Dro23k, Dro24a, Dro24d, Dro25b]. Interviews [Ano47w, Ano58h, Ano85d]. intimidate [Man15]. Intimidation [Par85, Sha85a]. Intolerable [Ano96k]. Intrigue [Pol01]. Intrigues [Boy00]. Introduction [Dro22c, Dro22b, Jr22b, Dro22d, Dro22e, Dro23f, Dro23g, Dro23h, Dro23i, Dro24f, Haf47, Has67, Hol96b, How65a, Küh17, Lew67c, Lew69a, LS59, McK25b, Mec12, Mec13a, Mec13b, Mec15a, Mec16d, Mec16e, Mec17a, Mec17c, Mec17b, Mec17d, Mec18b, Mec18a, Mec18c, Mec19d, Mec19e, Mec21e, Mec21g, Mec21f, Mec21h, Mec22c, MD24, Moo94i, Moo94j, Pri75b, Rab57d, Rab61f, Rab62b, Rab67b, Rab67f, Rab72b, Rin84b, SE58, Sno68c, Wil69a, Wil74a, WL21, Mec23d, Ack88a, Ber79, Moo95d, Thi46, Tsi82d, Tsi83c, Jan78, Lan82a, Pre58, Rab60g, Tsi77c, All54d, Ano89d]. Intuition [Ten75]. invaded [Lei89a]. invaders [Ano96l, Ark87c]. Invalid [LaF96]. Invasion [Dil57, Ad 90, Cot22, Isa90c]. Inventing [CK98, Hun04, Lan92b, Sin83, Pit91]. Inventor [Har92a, Kir63, Mor83b, WW62, Wil71]. Inventories [NA99e, KN13b, NK10b]. Inventory [Nor64, Rot93a]. Investigate [Rab49c]. Investigation [Ano49z, Met83, Mue69]. Investigations [Ano54n, Ano67k, Sta48, HIKWB10]. investigators [Kli80]. Investing [Ano97i, MD18, KK18a]. Investment [Nel57, Dro21b]. Investor [Cal64]. Invisible [Cot19, Moo99f, Gal13]. invites [Lev88]. Invoking [Kal53, Mel53b, Mel53a]. Involve [Ler49a]. Involvement [Bur74, Man66a]. Involving [Khr10]. Invulnerable [Ros75]. Ionizing [Ano56n, Lof84, Uni57, Coh79, Mor78, Sal90c, Tho12]. Iowa [Ano46-47]. IPPNW [Ano87-29, Bat86]. IQ [Alf81]. Iqbal [Das96]. Ira [Sti83]. Iran [AS16, AH92l, AH03b, AH03c, AH04b, Alb06, Ark95c, Arn98c, Ban11, Bin79, BF04, Brz87, Dal17, Gar08, HS13, Koc98, Kre87e, S.80]. Iran/arms [Kre87e]. Iranian [Alb95, Bar10b, Ben01a, Bur09, HM15, Kib10, Kor10, Lan10, Pic10, Ram10, SL13]. Iraq [Ram88b, Ack91c, AH91c, AH91d, AH91e, AH91m, AH92e, AH92f, AH92k, AH93a, AK95a, AO98, Ano91-38, Ano91-64, Arn90, Arn98c, Bac03, BWL21, Brz87, CRB^{+}09, CM91, Gro06a, Isa90l, Isa03g, Len21, Lov86, Moo91e, Pap98, Rub90, Sim91, Sob91b, Wur91b, Har92b]. Iraqi [AH91b, Fra21, Ne’81, Ram86a, Tay21, Tho80h]. Ireland [Opp02a]. Irenaus [Tho80g]. Iris [Kre98]. Irish [Fly98a, Opp03b]. Iron [Boy87, Yud83, Har86a]. Irony [dF79d]. Irradiated [Car55]. Irradiating [Ano69g]. Irradiation [AME50, Sch59c, War49, BL83c, LB78a, PGL91]. irrational [Bau84, Ema21, Moo97a]. irrationality [Ano87x, Shu87]. Irreconcilable [GG11b]. Irrelevant [SJ64, SK85, Gou77]. Irresistible [Bau84]. Irresponsibility [Rav65]. Irreversible [Fer62]. Irrigate [WC73]. Irvin [Haf79]. Irving [Nei82a, dF79f, Rab68g]. Isabella [Lev73]. Isaiah [Sha78a, Sha78b]. Isard [Kat53a, Rei63]. ISBN [Ano10a]. Ishikawa [Cog82]. Islam [Ahm04, Ano15g, Mus99b]. Islamic [And83, Hoo93]. Island [Ano80k, NM49, Sti83, Wal85b, Moo95j, Ano90-35, Cas02, Col79a, Fai79, Fel79l, Fri11, Gil80, HK13, KKHK79, Man04, SM14, Sti83, Wal83a, Doh57, Sti83, vH77a]. Islanders [Ham91c]. islands [Ung78a, Gom75a, Str56, Woo98c, Woo00d, Woo00e]. Ism [Dow80a]. Isn’t [TK66, Fel80d, Moo95p, Rof19]. Isolate [Kre97]. Isolation [Gol11]. Isolationism [Blo71]. Isolationist [Lin69]. Isolationists [Clo71]. Isotope [Ano47i, Ano48z, Ano49y, Day75n, Gar77c, Cas77e, PB84]. Isotopes [Aeb48, Ano46n, Ano48-38, Gle51, Kam46, Rob49, Kat54d]. Isotopic [Wol55]. Israel [Mos91, Ram88b, Ano87-35, Ben01b, Ben01c, Ben01d, Cal61b, CF87, CB06, Day92, Dos72, Fri74a, Gat72, Gel90, Git74, Kum95, Kum99, Lei90, Lev70, Lev72a, Mac77b, Maz81, Mei98, Mos88, Ne’81, Que69, Rab16, Sol90, SK85, Tur76, War92a, WP80, Wit01b]. Israeli [Lei92, KK72, KN14b, KK22b, Kum04, Lan14, NA90m, NAKH02b, Wei13, Wol93]. Israelson [Dav79b]. Issue [Ano49f, Bre68, Cam53, GRA^{+}76a, McM52, Tel60, Zei60, Zim48, Ano53q, Bey12b, Boy84, DM24a, Eps82a, Gas93b, Hea83, Kep82, Lei78, Lun82, Mor83a, PF82, Sch21, Sub77, You78a, Lin84a]. Issued [Ano46u, Ano57j]. Issues [Ano46-38, Ano52a, Aud65, Bas80, Fel72a, Fel81b, Gra81b, Hol77, Hol85b, Hum72, HM72, Joh65b, Lal64g, Lal64h, Lal64i, Lal64c, Lap72, Lov74b, Mor67, Mor49, Pre60, PvH74, Pri75b, Rod67, Sco72, Tam71c, Tam73a, Tri71, Ure54, Win84b, Ano92-34, Hol85a, Loc91b, Meh21, MS77c, Pug76, Rab56d, She85e, Smi82b, BGR61, Say84]. Istanbul [Fel73c]. Italian [Boe57a, Boe57b, Don57]. Italo [Fer69a]. Italy [BTCSC^{+}66]. Item [Lap51b]. Items [Ano54t]. ITER [Dro24a, Kat92]. Ithiel [Ano78p]. Itself [Ber64, Gre60, Mue07b, Woo99b, All18, Gou24, Tan83]. itty [Ano97n]. IV [Mil85c, Ano60d, Ano69p, Ano72h, Ark85f, BBH^{+}73, Bro56c, Bro86g, Coo56, Dan72, Lap55c, Lap72, Lev51a, Lew56, Men71, Orr56b, Wei56a]. Ivan [Ded60, Hag77, Ano94l]. Ivo [Bre87]. Ivory [Pei69, Tuc72, Wil76b]. Ivy [Lyn69]. IX [Ano53j]. Iyengar [Oze92].

J [Alc63, Alf84, All59l, And65, Ano46-107, Ano60f, Ano78f, Ano78n, Ano78l, Ano85d, Ano85b, Ano87l, Bal67, Bar73a, Bar90c, Bei84, BB63, Bou61, Boy87, Boy91, Bru57, Cah76, Cap98, Car68, Cha58b, Cra83b, Cra84e, Der88, Dow82d, Eri95, Fei77, Fel81a, Fel83e, Fer69a, Fra83, Fre84a, Gaf86, Gan82, Gar72, Gen84a, Gen84c, Gol55, Gol83c, Gol67a, Gol69b, Gro78a, Hal93a, Han92b, Haw56, Haw58a, HI72, How65b, dK60, Kat54a, Kin86, Kin69, Lap57, Law88, Leb94, Lew62a, Lof51, Mar79b, McC96, McK74, Mei68, Mil86b, Moo93a, Moo95l, Moo98b, Mul65, Ols62a, Opp62a, Par83, Pat81a, Per87b, Phi84a, Pla57a, PF75, Pre86, Rab73d, Ras63, Rei63, Rei78, Ros90a, S.80, Sch71a, Sco74, Sei69, Sei71b, Sei73, Sha50, She79, Shu78, Sin67a, Smi58e, Smi58f, Soc54]. J [Sto59b, Tel76, Ull63, Val73, WW62, Wea82, Wil84, You66a, You68a]. J. [Bet67, BS05, Bro05, GNMS54, GDH^{+}04, Kal54, Kal63b, McM05, Nic54b, Qua80, RB67, Uni54, WS79, Nor06]. Jack [Ano85c, BD79, Dow79c, Fel92, Lis58, Moo93b]. Jackass [Mac98]. Jackson [Tho80g, Men89, Wun82]. Jacob [Gar83a, Sto56a, Tay91]. Jacobs [Dow79c]. Jacquard [Mes64a]. Jacqueline [Ano91s, Bul08l]. Jacques [Lyn69, Nas82, Rab69b, She79, Mar53c]. Jacqz [Ano15d]. Jaczko [Sto19c]. Jahoda [Rea80]. Jahre [For80]. Jahren [For80]. jail [Day95, vH93]. Jaki [Fla68]. jam [Fos88]. Jamaica [Ven79]. jamboree [Wio77]. James [Alc63, Ano85d, Ano87s, Deu83b, Dow79d, Fer57b, Gen86, Gol95a, Gol78a, Har92a, Haw47, Her82b, Jac80b, Ker54, Lan98, Leb94, McC96, Mil69a, Mil85c, Mil85b, Pla57b, Rab68a, She82, SB91, Swe86, Van03, Wil53b, Wil71, You68a, Ano75e, Ano13d, Div70, Eav19, Jac73d, Lan90k, Rab64b, Vil75, Gol95a]. Jammer [Smi61]. Jamming [Wie06]. Jan [Ham91c, Lyt79, Tra71]. Jane [Ano78m, Sta85]. Janet [Bri86]. Janeway [Jan81a]. Janice [Kra94]. Janne [Tra90]. János [Van90b]. January [Ano64p, Khr63, vN57, Ano52s, Ano58d, Eps91, Rab56d, Tur52]. JANUS [Koe79, dF79b]. Japan [Ano82d, Kre89c, Van88, Ano57j, Ano80f, Bru47, BS01, Dow91, Fel82b, Fel14, Fra90b, Fri70a, Gom75b, GPR81, Ima69, Ima70, Ima71, Kan96, Kat16, Kaw78, Kuh66, Lev88, MNS46, NF13, Neu65, NA00d, NAB00, Oga58, Oga68, Ols62c, Opp62b, Pas55b, Pat88, Pow83, Sha87c, Suz11, TN90, Tak11, Tak17, TvH17, Tan88, Toy84a, Toy85, Ura01, Vog90b, Wei78b, Cap98, Has57, Kli90, vH77a, Dow86, Gin54, Ols62b]. Japanese [Fre95, Joh97a, Ker54, Mos81b, Sak69, Ano57h, Ano67n, Ano68h, Ele57, Kyo85, Kim87, Mor85c, Nis47, Sak88, SdSP62, YTS61, vH77a]. Jaro [Fra61a]. Jaruzelski [Kar88a]. Jasani [Lac85, Ano79a, Deu83b]. Jasanoff [Ahe91]. Jason [Eav17b, Bur74]. Jasper [SB91]. Jaspers [vKJ69, Spe68, Tor61]. Jastrow [Ano78n]. Jaw [Fel77f]. jaws [Isa91o, Nel92b]. Jay [Ano81f, Eli85, Keu83, Mei90, Moo95e, All54a, Ano57b]. Jayne [Dow83e]. Jean [Haf79, Rab69b, Rab73d, Tod81]. Jean-Claude [Haf79]. Jean-Francois [Rab73d]. Jean-Jacques [Rab69b]. Jean-Paul [Tod81]. Jeanne [LD79, Yar80a]. jeers [TM97]. Jeff [Ano89e, Els90, Dro23j]. Jeffrey [Ano78n, Ano86i, Ano87n, Div79, Lin84a, Mil85c, Mos90, Row88, dF79d]. Jehangir [Ano98n]. Jenkins [Sto18a]. Jennifer [Gil83, Mec20f, Bul09g]. Jensen [Jen70, Lew70f, Rab70b]. Jensenism [Bha70, PTNR70]. Jeopardy [Pea92b, Smy60, You56, Ano86i]. Jeremiads [Gre73a]. Jeremy [Sno68a, You66a]. jerks [Ano91-37]. Jerome [Tho80d, Cle95, Wou73]. Jerrold [Fel92, Leb91b]. Jerry [Lyn69, Val86, Ano14e]. Jersey [Dow84c]. Jerusalem [Bar80d]. Jervis [Har85b]. Jesse [Isa88f]. Jesus [Rab73d]. jet [Fly97, Har90d, Ste95]. jets [Oel20]. jettison [Rob20]. Jews [Rab75, Sma76a]. JFK [Hig76]. Jiankui [BV20]. Jibran [Gau24]. jigsaw [Bey12a]. Jihad [Lei07, Mus00]. Jimmy [Phy80, Sig98, Wil78a]. Jo [Rep86, Tay98]. Joan [Dow82d, For80]. Joaquín [Ano91u]. Job [BA07, CP99, Moo00b, Moo00c, Bal89a, Row77]. Jobs [Lum94b, Rab69g, Lum94a, NA88k, NA89i, Rot93f, Zam92b]. Jodrell [Coh69]. Joe [Dow80a, Ano15e, HK21]. Joel [Day75c, Gol93a, Mac84b, Mor79, Sil79a, Yon68]. Joffe [Sim88]. Johansen [Woo82]. Johanson [Met83]. John [Ack95, Alf84, Ano50-41, Ano78n, Ano85d, Ano85b, Ano86i, Ano90m, Ano90o, Ber66a, Ber77b, Boy89b, Bro47c, Cra83b, Day75b, Day77b, Eck78, Eis67, Fer57a, Fol74, For87, Fre76a, Fri70b, G.84a, Gar97a, Gol96a, H.84b, Hal93a, Hul98, Lac85, Ler80a, Lyn78, Mar78a, May89, McK74, Mye87a, Nic54a, Nor04, Pal64b, Rad83, Roc81, Sch90, Sha81a, She83b, Sie85, Sil71, Sno69, The89e, Ten75, Uhl98a, Wea82, Wil92a, Wou73, Wur68, You84b, vL92, All58a, Ano63q, Ano68g, Ano87-32, Ben07, Dow80e, Edi60, Moo00g, Bul08m, Tel57a]. Johnson [Ano45k, Bri86, Dow83e, Jac80b, Mil85b, Rau91]. Joie [Hul76]. Join [Bar10a, Zhe90c, Ano85y]. joined [Gan83e]. Joining [Pow98]. Joins [Kha98]. Joint [Ano48-36, Ano53w, Ano54h, Bla69b, RDS^{+}52, Zag94a, Ano49z, Ano49-36, Ano51r, Ano53a, Ano53b, Ano53c, Ano53d, Ano53w, HPL^{+}49, Jac74e, Mat98e]. joke [Som84a]. Joliot [Rab58b, Alb54, Ano49-27, Ble54, BL81, CCD^{+}52, Hai54, Mer54a, Rab50d, Rab54c, Rum54, Tod55]. Joliot-Curie [Ano49-27, CCD^{+}52, Rab50d, Rab54c]. Jolly [Ano96n]. Jon [Ada87, Mec19f]. Jonas [Ehr82b, Gus85]. Jonathan [Ano98h, Bir87, Ehr82b, Har88c, LP92, Sar85, Sha95b]. Jones [Dow82e, Smi58f, Bre74, Cra83a, Fis83, Smi58e]. Jong [Mye88]. Jordan [GJ67]. Josef [Mul66, Pol68]. Joseph [And71, Ano57f, Ano85d, Ano98h, Cra84c, Ham61b, Jac80b, LaF62, Moo93a, Moo93b, Nat87, Pee78, Per87b, She79, Shu78, Wal85a, Ano47w, Ano92-32, BM99, Cor94, Lan96, McK25a]. Josephine [Ano80l]. Joshua [Ano86j, dF79f]. Joughlin [Din69]. journal [Ano89b, Ano89w, Moo91c, Zag91a]. Journalism [Ano91u, Dor84, Les63, Col84c, Dro23m, Gou24, Nus79, Pan86a]. Journalist [McC88a]. Journalistic [Gar85]. journalists [Wei91a, Zag91b]. Journals [And65, Gre67b]. Journey [Fol74, Pal02, Spi64, TS01, Van03, Wei54a, Eng91, Gla12, Mal79, MK02a, Wal21a]. Jouvenel [Rit67, GF64]. Joyce [Mye89, Com48b, The48]. Joyride [Fre73]. Jozef [Ano88e, Don86b]. JPL [DN67, Nie66b]. Jr [Ano50-41, Ano85v, Ano89f, Bri86, Fin90, Jac80b, Jas81, Met83, Pon67a, Saw79, Sta64, Wal81a, You66c]. Jr. [Ano74f, Ber73c, Cra84c, Ehr81, Eng91, Fie24, Gen86, Gol78a, Gre58, Jac74c, Kru74, Lap54c, McC93, Nat87, Nie63, Rob69, Rus51a, Sei71b, Swe86, Wal85a, Win81a, Woo65, Woo73b, You68a]. Judge [Asp80, Röl81, Eat78]. Judging [Har10]. Judgments [Ing68a]. Judiciary [Her88b]. Judith [Dow81a, Pal02]. Jules [Ano52y]. Julian [Dal82, EE85, Mor59a]. Julius [Mon53a]. July [Ano55p, Ano55q, Ano59d, Che52, CD48, Joh68, Rab56h, Ano68f, Gro70, NCA88a, NA88d, NA89e, NA90b, NA91d, NA92d, NA93d, NA94c, NA94g, NA95e, NA96e, NA97d, NA98f, Rab54g]. jumped [vMC98]. Jumping [Rot96g, Tor77a]. Jumps [Isa03f]. Junction [Lew68f]. June [ADH^{+}73, Ano46-90, Ano59d, Jac71b, Lie46, SZC54, Tur52, Ano58i, Ano81g, Ano81h, Ano87-45, Car81, NA88t, NA89q, NA90r, NA91p, NA92p, NA93q]. Jungk [Bet58, Liv69, Rob62]. Jungle [Wil75e, Pol91a]. Junk [Kra06]. junkies [Ark93a, vHZC93]. junkyard [JR89]. Jupiter [Jac73d]. Juraguá [Fly98c]. Jurassic [Ano96m]. Jürgen [Cre68]. jurors [BM93]. Jury [AO95b]. Jury-rigged [AO95b]. Just [Ade91, Ano47l, Ano89s, Ano90-32, Ano90-36, Ano91-36, Ano91-41, Ano93r, Ano94n, Aue03c, Cai90, Coo06, Den82, Dom97, Gro98, Isa93o, JM85, Kaf98, Lym96, Mil85b, Moo95c, NA90d, NA91e, Tay92, Vog90a, Web03, Weg96a, Ano84l, Ano98n, Bli76, Fie21b, HPR91, Kib12, Nay12, Nov77b, Pat80b, Rot94i, Squ12, Wal21b, Chi78]. Justice [BSS06, FJ91, Koc04, Baz83, Dan54a, Gan76, Koe22, Ano88n]. Justification [Rin75]. Justified [Bla67, Hol94c]. Justify [JL74, Bea86b].

K. [Bul07d]. Kabul [Win90]. Kafka [Moo94b, Zag94f]. Kafkaesque [Moo99b]. Kagarlitsky [Dar91]. Kahin [Wur68]. Kahn [How65a, Mei68, BA07, EE85, Fri62, Kea61, Pet61, Por82, Sto61]. Kai [Dro23k]. Kaku [Gil83]. Kalashnikov [Kla99]. Kaldor [Kre83a, Per91a]. Kalinga [Aur71c]. Kaliningrad [Woo03]. Kalven [Ano74f]. Kamen [Gen86]. Kansas [Ano46-48, Ano71g, Ben79a, Bor46]. Kant [Elm71]. Kanter [Sco79]. Kanti [Das96]. Kantrowitz [Kan71b]. Kapitsa [Fel84f, Ano90-61, Kap90a, Rot93v, The90, dW90a]. Kapitza [Tuc62]. Kaplan [Cra83b, Roc81, Rod85]. Kappen [Ano89g]. Karaska [Rei63]. Karen [H.84b]. Kargil [Tar23]. Kargon [Bri80b]. Kariba [Scu65]. Karkoszka [Sar78]. Karl [Spe68, Tor61, vKJ69, Tur54]. Karlesky [Shu78]. Karlovy [Ano64q]. Karlson [Sch69a]. Karman [Gre68b, Gre68b]. Karpin [Van03]. Kartiganer [Lad81]. Kashmir [Gan03, Fai97, Obe97a, Obe98, Sha99]. Kasperson [LD79, LD79]. Kastler [Ano84b]. Kate [Ano98j]. Kates [Rei78]. Katharine [Ano13e]. Katherine [Pre96]. Kathleen [All54b, Bro55b]. Katmandu [Gom75a]. Katrina [Ano90o]. Katz [Rom82]. Kauai [Jon93a]. Kaufman [Dow80e, Rep86, Sch80c]. Kaufmann [Pap88]. Kazakh [Ust96]. Kazakhstan [Bel02, Zag93c]. Kearney [Smi83]. Keep [Ark89a, Blo66, BF05a, Cae74, For92, Gya90, Kis89, Ler46, Mot49, Nit92, RP66, Ros75, Spe89a, Wil75b, Ald76, Bai95, Dro25c, Isa89m, Isa90h, Kim20, Mec19c, Rot95j, Zha12, For20]. Keepers [vL92]. Keeping [Ano93u, BST02, Cha89c, Fin65, Gol61c, Mad63, Opp60, Rot93l, Yar80b, Yar80c, Cox86, Div70, Fin65, Mar85a]. keeps [Eps91]. Kefauver [Ano55e].

Keg [Dje75]. Kehler [Ano83f]. Keisling [Sti83]. Keith [Che88, Hob85, McG84, Wun82]. Keithly [Ano85c]. Kelber [Ano89f]. Kelleher [Sco74]. Kellner [Hal93a]. Kellogg [Bal79c]. Kelly [Ano53t]. Kemeny [Gre80, Lan80]. Kemp [Lyt79]. Kendall [Dow81c, Moo99e]. Kennan [Gus83, Ken87a, Lyn69, Sim54a, Iri86]. Kennedy [Ano63q, Smi85d, Ano62w, Bro62a, Edi60, Ken60a, Ken61a, Ken61c, Ken63b, KK63, Ken71, Lew68e, Szi61a, Lon82]. Kenneth [Ano57b, Eis67, Got64, Har92b, Kre90, M.85, Was62]. Kennett [She83a]. Kenton [Ano85d]. Kenya [Odh79]. Kept [Mor47b]. Kerry [Isa04a, Sto17b]. Kessle [Tra71]. Ketchum [Fin90]. Kevin [Dow83e, Uhl96, Eav18]. Kevles [Mar79a]. Key [Ano96x, Aue06c, Bro57a, Sto68b, Wad58, BHK^{+}21, Isa86a, Sig85a]. Keynesianism [Som91]. Keys [vL92]. KGB [Bou98b, Gev93]. Khan [Mus01, AH92g, Men92b]. Khariton [KS93]. Khasbulatov [Chu93]. Khrushchev [LaF86, Leb91b, Lon82, Ada95, Ano59e, Ano64m, Gar88c, Her60, KK63, Khr57, Khr63, Leb88b, Leb91a, Leb91b, You66c]. Kick [Uhl97a, Mal65]. kicking [Gro98]. Kickstart [Kib12]. Kidder [Dro17]. Kiev [Kis93c]. Kill [ARA94a, Com74, Dro23j, McL65, Ter64, Aft94b, Ham91b, Kul88, Rot94h, Kal63a]. Killed [Kal95a]. Killer [Aue03d, Gub14]. killers [Ano93-37]. Killian [Gen86, Gol78a, Wyl56]. Killing [Dow80j, Hay76, Lor70, Mar06a, Moo99d, Woo97c, Lif13, NA89i, Wal90d, Wur94c]. kills [Bel93]. Kilowatt [Gol90b]. Kilowatts [Dow83c]. Kim [Dow82g, HC12]. Kimball [Hig66, Kat52a]. Kincade [Div79]. Kind [Lon99, Mus99a, Ric74, Bus94, Jr22c]. Kinda [Isa97a]. kindness [Loc89a]. King [Ano83b, Dow83f, Abb89, Gut93a, Ano45q]. kingdom [Ano93x, Ano60c, Bar82a, Brz12, Buc60, Fre81, Kal81, KK21e, KKJK24d, Low85, Mot83]. Kingdon [Haw58b]. Kings [Chr71, Ell56, Woo97f]. Kinoshita [Dow83f]. Kintner [Ste76a, Lap55a]. Kirgizstan [And94]. Kirincich [Dro21c]. Kissinger [Mul65, Day77h, You66a]. Kistiakowsky [Ano85a, Rat83, Hor77]. Kit [Fow59]. Kitzbuhel [Ano58l, Ano58k]. Kitzbühel-Vienna [Ano58k]. Kiwi [Cle92]. Klare [Har86e, Moo95b]. Klaus [Ano50r, Ano57d, Lal67a, Sei60, Che52, HM90b]. Klaw [Wil69c]. Klein [Dow82d]. Klochko [Rab65c]. Knebel [Wil68b, Mor61a]. Knee [Wei89a]. Knees [Ste89f]. Knew [Gla75a, BV20, Cri94, Hib96a, OM97]. Knight [McA71]. Knightley [Wei89b]. knock [Tov89]. Knockoff [Har05]. Knoll [Lue96, Day97]. Knopf [Ano87n, Ano89e]. Knorr [Lal67a, Ano57d, Sei60]. Knot [Men92b]. Knots [Mei06]. Know [Ano95n, Gar97a, Gen84c, Gol98, Nad04, NAB00, Rab54b, Smi54c, And93, Ano93y, Lan89e, Mac11, McK25b, MHM90, NA90l, NA00d, Nuc20, Sie24, Wie57, WA93, Zwe84, Car53]. know-how [NA90l]. Knowing [Ano96j]. Knowledge [Gol96a, Hir65, Hol65, Hur78a, Sno69, Ste80b, Tan83, Zil64, Kri14, Wyl53, Mac97a, Ols59]. Known [McM48b, NCA88a, NA88e, NA89h, NA89g, NA90e, NA91f, NA92g, NA93h, NA94b, NA95c, NA96c, NA98c, NA90m]. Knows [Lee60, Isa89m]. Koestler [Sto59a, dF79b]. Kofmehl [Got64]. Kohl [RK89b]. Köhler [Lyt79]. Kokoshin [Bou87]. Kola [RMS03]. Kolbert [Ano14b, Dro24e]. Kolko [Mer90]. Kolkowicz [Eva86]. Koller [Kat54e]. Komarov [Ehr81]. Kondrashev [Bou98b]. Kong [Gal91, Gom75b, Hum83, Ste97]. Konrad [Rab66b]. Korea [Ahn10, ABKS10, AH91n, AH92h, Alb93e, Alb94b, Alb94d, AH02, Alv03, Alv20, Bin10, Cum92b, Cum15, Gil98, Gol09, GP18, Har90d, HSV19, Hay93, Hec08, HC12, Hym07, KHB^{+}05, Kan09, Kan11, KKL15, Kim20, KN17a, Lan16, Mit50, iM16, NA87e, NKH03b, NK05d, Pan22, Rab51k, San91, She16, Sig01, Sig10, Bul03b, Tug51, Wat19, Woo98b, Wri99a]. Korean [Zel92, Ano92-33, Cha11, Har90b, Kim95, KN18b, KK21b, KK22c, KKJK24b, LST10, Moo96l, PPE22, Rab50b, Sam18]. Koreas [Cum98]. Kornai [Van90b]. Kosta [Cal84a]. Koughan [Dow81g]. Kovel [Mac84b]. Kovner [Kol63]. Kowarski [Fel79o]. Kraatz [Len21]. Kraatz-Wadsack [Len21]. Kramish [Bro57a]. Krane [Bar90b, Dow84a]. Kranzberg [Saw81a]. Krasnoyarsk [DCM87]. Krebiozen [Ivy64, Mar64g]. Kremlin [Tum87, Gri87a, IM97, Kor89, Nea60, Que81]. Krepon [Her85]. Kriegführung [Ano68l]. Kriegsaktivitätkeit [vL48a]. Kriegstätigkeit [vL47]. Kristensen [Ano14a]. Krosney [And83]. Kruglak [Sal63b]. Krupp [Sea69]. Kruzel [She79]. Kudos [Don97]. Kuhn [Nor64]. Kula [Ora65]. Kull [Ede89]. Kummer [Rab73c]. Kurchatov [Hol86a, Rab67f]. Kurds [McK99]. Kurt [Bri86]. Kuwait [Ack90d, Isa90c]. Kuznets [Str67]. Kuznick [Der88]. Kwajalein [Woo00e]. Kwitney [Har88c]. Kyoto [Ano57j, PR08]. Kyrgyzstan [Dei05]. Kyshtym [Bru90].

L [All56l, Ano45n, Ano50-41, Ano78d, Ano78n, Ano87n, Ano89e, Bal79c, Cha74, Cog82, Cro67, Dav69, EE85, Fel73a, Fel81a, Fel84f, Fla68, Gal86, Ham61b, Hun85a, Kin80b, Kre90, Lac85, Leb91b, Lyn69, Mar78a, Mun54, Nel84, Pal02, Pla57b, Rus51a, Sch68e, Sha50, She89, Shu78, Smi83, Spi66a, Sta64, Ste86c, Sti83, Swe86, Tho80g, Uhl95a, Van03, You66c]. Lab [Alv74, Rez07c, Ano90s, Ano90-39, Ano91x]. label [Sin92, Vog90b]. Labor [Ano46-49, Pie59, Pos21, Gap90, Mos89]. Laboratories [Cle48a, DuB46a, Gre72, Mor70a, Mor70b, Par66c, Spi66b, Spi67, Tel50, Ber84d, Das85, Rep10, Ano46-28, Duf70, Smy50b, Smy64]. Laboratory [Ano48d, Ano53e, Cra70a, Cra64, Dan70, Der88, Mit70, Moo97a, Wen53, Bru76, Ano46-73, Bra54, Dan48, Eis54a, Fer54a, Gus11, Man49a, Man49b, Mor46, Sta00, WG69, Yor75a, Tay98]. Labors [Rab51f]. Labour [Gap86]. Labs [BCMR92, Cho02, DeW84, Man74, Arc86, Bal86b, Bev86, But97b, Day77c, Day77k, Den14, Kah86, May85, Sea89a, Sea89b, Tur13, Zam92b, Zam93]. Lac [Ano58g]. Lack [Bor85, Gué69, Ker76, Zif53]. Lackey [Mil85b]. Lackluster [Woo01b]. Ladder [Gou75b]. Laden [ABH02]. L’Affaire [Ano54p]. Laffin [Ano90m]. lag [Led85]. Lagemann [Whe84b]. Lagoon [Fly00c]. Laid [Rom92]. Lake [Got69]. Lakes [Scu65]. Lakoff [Lew62a]. Lall [Ano78o, Wri67]. Lament [Day75i]. Lamont [Fel66b]. Lance [Dow83e, NA87e]. Lancelot [Cro67, Lan76, Smi59b]. Land [Ano46g, Axe51k, Buc50, Gou70, Jac73b, Lap55a, Moo97l, Rab98, Vof72, Wur94c, Wur96c, Wur96b, Col79c, Rot94g, Wol17, Wur93b, Wur93c]. land-based [Wol17]. Landa [Bel93]. Lander [Jac71a]. Landing [Har69b, Lew69a, Wat73a, Buk93, Cav92, Lov66]. Landings [Hin69b, O’K69]. Landless [Mus70]. landmine [Rot98h]. Landmines [Fly05b]. Lands [WC73]. Landscape [Dub70, MM11, Nac69, Gol24]. Landscapes [Fra81]. Lang [Tel49, Bro55a, Bro60a, Tel49]. Langdon [Aue78, O’C87]. Langer [Bla84]. Langguth [Swe86]. Langley [Pra93a]. Langsdorf [Ano96f]. Language [Ano87s, Bru72, Hob78, Tam71a, BB87, Coh87, Dro23b, Her60, SHT^{+}87]. Languages [Cha72, She85d]. Lanouette [Low93]. Lansing [Fel66b]. Laos [Lov00]. Lapp [Mar58a, Fin57b, Kal63a, Lis58, Rab53a, Shu49]. Lapses [Ano49s]. Laqueur [Swe86]. Large [Coh69, DuB46a, Kow49, McC57, Mer24, Ran55a, Zag94c, Zag94b, Zag94a, Zag94f]. Large-Scale [Kow49]. largely [GG16]. Larger [Ber04b, Bar77b]. largesse [Ano91-30]. Larish [Dow80i]. Larry [Bal79c, Dow80e, Sha81a, Tho80d]. Lars [Sla88]. Larsen [Smi54a]. Larus [Yon68]. Lary [Rit67]. Lasby [Sea72b]. Laser [BF05b, Cas77b, For02, Gar77c, Gho11, GC77, PB84, Pri03, SB80, Ano86-32, Ano87-40, Blu88, Cas77e, DeW88, Dic77]. Lasers [Fri88, Kap83a, Shi88, Wur96a, Ark95d]. Last [Cha75c, CO06, Dow81h, Dug91, Fly06a, Haa78, Han00a, Khr00, NA87g, NA92k, Pai82g, Pet50, Rab56j, Ste89c, Ure48d, Yor69, AK95a, Ano93y, Ark94a, Ark94d, Ark94b, Ark94c, Ark95a, Ark95c, Ark95d, Ark95b, Ark95e, Ark96d, Ark96a, Ark96b, Ark96c, Ark96e, Ark97c, Ark97b, Ark97f, Ark97d, Ark97e, Ark97g, Ark98a, Ark98c, Ark98b, Ark98d, Ark98e, Ark98f, Gom82a, Isa88f, Jac92, Lan89e, Loc91b, Mye82a, NA91h, Pom79, Rao87, Sut96b, Wee98, Wil92c, Zal19, Dow82h]. last-ditch [Wil92c]. Laszlo [Mye88]. Late [Myr68, Pri03, You68a, Gei16a, Lun79]. lately [Ano94m]. Later [Ano55m, Ano66f, Fra54, Jac73f, Lal67e, Lap53, Mil78, Rab70e, Str06d, Ano97i, Ber84b, Jr16, Gor11, Gor20, Hol95, Pai83b, Rot81b, WB95, Zag95e, H.85]. Latest [Hor88a, Lon87, NA91g, SH94, The88]. Latham [Sta64]. Latiff [Mec22a]. Latin [Bar67b, Mye87a, S.80, Sla88, Ast67, Bar66b, Bon72, Cli77, Fal70, Fly97, GS81, Gol82a, Lop66, Str81]. Laue [For80, Mor48a]. laugh [Haa78]. Laughing [Wie78]. launch [Ano87-38, Leb87]. Launched [Rin75, Ano88p]. launches [Ano91-52]. laundering [Arn22]. laundry [Aft94a]. Laura [Cam54, Hug57, Gro78b]. laureate [Fra76a, Kas75, Mec20f, Nel77, Tis52]. Laureates [Nad79, Day77i, Ang78, Bal79a, Bun79, NB79, Orr78, Pau79, Sta79, vS78]. Laurence [Ber77b, Roc81, WW62]. Laurie [Ano12d]. laut [Ano75g]. Lavakare [Ano78l]. Lavon [Wei13]. Law [Dyc84, Fra61a, Fre88, Fri82, Gre67c, Gre63b, Jas81, Jon63b, Kat58, Lan76, Liv68, McK74, Men91d, Mil48b, Pon67b, Sch62a, Smi59b, Tha78, Way51, Yut69b, Arb85, BW18, Fal89, GG11b, Gre77, Gre23, Gre12, HS97, Hud77, Hun85c, MA17, Pol91a, Sol91, Ste91c, Wei18, And65, Ano64j, Kuy54, Nad98d, Pat75, PP63, Rev83, Sha86, Sha59, Swi83, You78a, dK60, Sei73, Bre74, Dam62]. Lawren [Gol88b]. Lawrence [Div79, Hoe88, Mor85a, Mos81a, Olu71, Roc83a, Wal87a, Win94, Win98, Ano52k, Dav68, Dav86b, Sto24, Yor75a, Wil69b]. Laws [GF73, Lan57, All14a, YES^{+}21, Par62]. Lawyers [Kat58, Lap57]. lay [Car97, DM18]. lay/expert [DM18]. layered [CS17]. Laying [Hin18, RS88c, Twi23, Mat98b]. layoffs [NA88h]. Lays [NA92n, NA92h]. Lead [Bro55a, Sch54c, Tha00, KGS25]. Leader [Com75d, Cor53, Lev64, Lum99, Hol91, Lan90c]. Leaders [Ano46-49, Ano46-59, Deu68, Hes85, Ran82, Tru85]. leadership [Kre11, Pon17]. leads [Ano90-42, Par85]. Leaf [Hod00]. league [Smi90c, Ano46-50]. Leaks [Sch00a, Bol88a, PR14, Pow86]. leaky [Mül93]. Leap [BCMR92, Nee79]. Learn [Ano54-30, Dro23e, Wu23]. Learned [Boy05, Gué67, Kin86, Ber23, Coh87, Jr16, Fri20, FD23, Gor11, Gor20, Kat16, Mar77, Tin23, ZA04]. Learning [Bro79d, Bro79e, Isa89k, Mat80b, PPE22, Smi80]. Lease [All60h, Ano91-36]. Lease-a-MiG [Ano91-36]. Leasing [Agn82a]. Least [Lee60, Ano93r, Sch24]. Leave [Ark97d, LC88, Mur54c, Ano92w, Ano94k, Sho89]. Leaving [Rot85, Far13]. Lebanon [Kal78]. Leben [For80]. Lebow [Kra94, Bla84, Ble87, You82]. LeCron [Ano86i]. Lecture [Saw81b]. Lectures [Pla65]. Led [JvH09, Wei13]. Lederberg [dF79f, Led67]. Lederman [Fla90, Ano93-28]. Lee [Ano87-32, Fin65, Gre68b, Lac85, Sch68e, She83a, Sto59b, Wil80b, MJ68, Par86]. Lefever [Pom66]. Leffler [Per94]. Left [Ano98q, Bor64, Llu00a, Mei51b, Mor00, Ber79, Dar91, Llu00b, Lyn69, Pol91b]. leftover [Var96b]. Leg [Khr00]. Legacies [Str07b, SB23]. Legacy [And74b, Ano80k, CRB^{+}09, Com75e, Dan90b, DeW62, Gov93, Gra93, HP09, LT99a, Mas99b, MG08, Mor62b, PMS00, Ste88, DM24c, Fel81f, LT99b, Mak98, Pra22, Rab98, The89f, Tho79b, Tud21, WH11, Wee98, Cul08, Dow82e]. Legal [Ano78d, Bur50, CB95, Cav53, D’A70, FCC^{+}46, IK21, Kal60, Kal53, Kri85, Lap57, Mel53b, Mou73, Alm85, Gre81, Kal83b, Lef77a, Mey83, Roc18, Sei71b]. legalistic [Sta80]. legalities [Far13, RH12]. Legend [Shi51]. Legends [Dra06]. Leger [Ano90p, Goo86, Rep83]. Legionnaires [Gol95c]. Legislation [Ano46-36, Ano46-96, Ano47a, DeC48, May47a, Par46b, Par47, Ste46, Ben87, Mec21c]. Legislators [Dys75]. Leitenberg [Ano88c]. LeMay [Lap69]. lemons [RM98]. Len [Hig66]. Lend [All60h]. Lend-Lease [All60h]. lengthening [Mar93a]. Lennon [Lag81]. Lens [Day77b]. Lentz [Sto56b]. Leo [Gab73, Hut61, Rab64b, Ano79f, Ano79g, Ano79i, Ano79m, Ano94q, Bes85, LS94]. Leon [Fel73a, Fla90, Kin86, Mac62, Sha84b, Van88]. Leona [Wil80a]. Leonard [All56k, Ano78n, Ano91s, Don86b, Kel98a, Kin80c, McE88, Nor64, Rab64a, Sac72b, Tuc97, You66a, Moo99g]. Leonid [Phy80]. Leonidovich [The90]. Leopold [Ber79]. Leptons [Ano78m]. Lerman [You68a]. Lerner [Rab72a]. Leslie [Gol88b, Lei92, Nor02b, Nor03, dF79d, Ber03]. Less [Ano95k, Buk02, CZ95, Etz07, Isa89i, Kre92a, Lan96, Ric01a, Bul03b, Sto96, Tel76, Ure49a, Dro20b, Eav18, Isa97c, Moo95n, Off84, RS16, Rot81a, Ano92-30]. Lesser [Gla75c, NA90f, NA90g, Ehr82c]. Lesson [Lap55f, Nac72, Rab69d, Tuc97, Agn77a, vCM89b, Wal86]. Lessons [Bla85b, Cam88, Cir05a, Day75k, Fed59c, Fel85b, FD23, FWCL20, Gol70c, Gol11, Gué67, Hod81b, Hol85b, Huq11, Kin86, Kon67, Kre11, Lew63a, Man23, Myn11, Pet50, Rab56c, Rab63b, Rev62, Ric12, Roc78, Sha10, SW21, Tay21, Tin23, WKK23, Gor11, Gor20, Kat16, NA88b, Suz11, Jac92]. Lest [Day77i]. Lester [Ano78n, Bal79b, Ehr83, Han75, Rab73e, Roc85, Tod79, Wou81, Ano14f]. Lestingi [Dow82f]. Let [Ahm00a, ABB^{+}50, Ano94m, Ano95n, Ano98l, Car94a, Col74, Com50, Dah77, Day75f, Fel55, Fel77g, Fel80e, Ing59b, Isa89k, Mak94b, MvH97, Obu89, Pom65, Rab47e, Rab62k, Rei93, Roe91, Sha95c, Siv75, Tho77, TS92, Wei68b, Zim87, Ano92s, Flo88, CP99, PS04]. Lethal [Ack90f, GT06, Rab98, Aft94a, Aue03d, Cal17, Ano94s, Ros94]. lethargy [NA88f]. Letter [Abl85, ATL85, Abr76, AG63, AC76, Ago77, Alb54, Alb57, Alb64, Ale60, All86, All88, All65, All57k, Ama77, Amd65, Ame66, Ams76, And58, AL62, And78b, ADH^{+}73, AL79a, Ano48x, Ano62v, Ano62w, Ano67j, Ano80m, Ano86-35, Ano02k, AH62, Apt64, Arm53, Arn54a, Arn98b, Arp61, Art85, Ash60, Ash76, Asi57, Asi62b, Aud65, AC86, Aug56a, Aur71b, Aur71c, Aur71a, Aur71d, Aur71e, Aur71f, Aur71g, Aur72b, Aur72d, Aur72c, Aur72a, Aur72e, Aur72f, Aus60, BJCP88, Bac77b, Bae60, Bak86, Bal93, Bal87, Bal76, Bal89a, Bal60b, Bal86b, Bar61b, Bar65, Bar74, Bar64b, Bar85a, Bar78c, BP85, Bas86, Bat52, Bat86, BPR52, BD76, Bea87, Bea86b, Bec62, Bec78, BR85, Ben61, Ben78, Ben79b, Ber76a, Ber77a, Ber87a, Ber86c, Ber66b, Ber78b, Bes94, BE62, BS67, Bet60, Bet89b]. Letter [BCMR92, Bet93, Bet95, Bev86, Bey85, Bey46, Bir76, Bla87b, Bla88, Ble54, Blo66, Blo95, BCW74, Boe52, Boe76, Boe57b, BA89, Bon90, Bor85, BR57, Bos78, Bos98, Bou77, Boy86b, Bra95, Bre85, Bro89a, Bro76a, Bro66, Bro93, Bro47e, BGR60, BH65, BW77, Bro87b, Bru96, Bru88, Buc77, BS85, BT95, Bur90a, Bur90b, Bur48, Bur50, Bur88, BTCSC^{+}66, Bye87, Bye88, Bye91, Cab97, Cae74, Cal60a, Cal60b, Cal61b, Cal86a, CA98a, Cal57, Cam77, Cam53, Cam60, Cam76, Cam88, Cam90, Cam95a, Cam95b, Car77, Cas85a, Cas85b, Cat87, Cat89, Cha64, Cha77, Che89, Chi59b, Cla90, Cla54, Cla59, CO74, Cob78, Cob88, Coe98, Coh60, Coh62, CS63, Coh76b, Coh76a, CD77, CEE85, Coh95a, Col95, Com76a, Con60, CE88, Coo58, Coo76b, Coo76c, Coo77, CMMF62]. Letter [CO85, Cox86, Cra98, Cra76, Cra61, Cun88, Cut58, Daf64, Dar86, Das85, Dav57a, Dav74, Dav55, Dav65, Dav76a, DFJ^{+}63, Dav77b, Dav77c, Day93, Day95, DeM76, DeN65, DeN76b, DeN76a, DeN78, DeV88, DeV92, DeV62, DeW88, Den47, Den77, Der85, Die91, DR92, Dix66, DFF^{+}63, Dol76, Dom97, Don57, Don64, Don85, Don88, Don97, Doo64, Doo60, Dor96, Dou60, Dow92, Doy53, Dri93, Dua78, Dub57a, Dud77a, Dud77b, Dum75, Dur69, Dys59, Edm77, Eds62, Ein49, Ein53, Eis90, Ell76, Elm66, Elm76, Elw64, Eme54, EG86, Eps77a, Eps88a, EP86, EG89, Ewi60, Fai97, FP89, Far98, Far86, Fau87, Fel62b, Fel66c, Fel85b, Fel88, Fel66d, Fel86b, Fel57, Fer54a, Fer57c, Fer60, Fer87, Fes76, Fes77, Fie78, Fin57c, Fin92, Fis76, Fis89, Fla60, Fle59, Fle58]. Letter [Flo85, Fol60, FW64, FM93, For70, For64, Fot89, Fow55, Fow58, Fow59, Fox52, Fra65, Fra90b, Fre57, Fre76c, Fre77a, Fre78, FC96, Fri78, Fry58, FD76, Gab86, Gap86, Gap90, Gar85, Gar74, Gar77c, Gar86, Gar97b, Gel90, Ger58, GF64, Gil77, Gil65, Gil98, Gin76, Gla75b, Gla76b, Gla76d, Gla76c, GC77, Gla63, Gla78, Gle52, Glu50, Gne77, Gol93b, Gol88c, Gol88d, Gol88e, Gol90a, Gol97, Gol73a, Gol73b, Gol77, Gol89, Gom78, Gom62a, Gom62b, Gom70, GRA^{+}76a, Gom87, Goo77, Gor85, GG63, Gou59, Gou61, Gou63, Gou78a, Gou95, Gou75a, Gra76b, Gra61, Gre47, GC63, Gre63b, GW65, Gri87b, Gru78, Gui77, Gyo77, Haa77, Haa78, HKC78, Had77, Hah61, Hai86, Hai54, Hal76b, Hal76a, HWG95, Han89, Han90d, Han92c, Han95b, Han96b, Han97]. Letter [Han98b, Har61, Har88b, Har91a, Har59, HA85, Har77a, Har60b, Har64b, Har65, Har62, Har77b, Has85, Hat85, Hau89, Haw94, Hay55, Hay76, Heb52, Hec78, HH65, Hei89, HJ76, Hel85, Hel77a, Hel77b, Hep77, HI94, Her95b, Her78a, Hes97, Hes55, Hig76, Hig63, Hig89, Hin60, Hir89a, Hir89b, HM97, Hla65, Hob78, Hoc57, HR51, Hog86, Hoh75, HWH78, HW96, Hol04, Hol94c, Hoo63, Hor60, Hor76, HJS^{+}92, Hot76, Hou76, How63, Hub57, Hub60, Hud60, Hue77, Hug87, Hug95, Hun90, Hun85c, Hur74, Hur77a, Hur77b, Hur78b, Hur88, HE76, Hyd91, Ibs87, Ich64, Ide93, Ing64, Ing77, Ise97, Ivy64, Jag95, Jag98, Jan81a, Jan76a, Jan76b, Jas91, Jau90b, Jef94, Jeh60a, Jeh60b, JvL63, Joh64a, JM85, JC54, Jon63a, Jon90a, Jon91b, Jon93b, Jon97]. Letter [Jor67, JR62, Kaf98, Kal77b, Kal85b, Kal95b, Kal85c, Kap89, Kap90b, Kap93, Kar97, Kas60, Kat95, Kat54b, Kaz61, Kea60, Kee56, Kee76, Kel62, Kel87, Kel90, Kel98b, Kem97, Ken87c, Ken77, Kep76, Ker76, Key57, Kim95, KH63, Kin67, Kin74, Kir55, Kir91, Kis54, KA88, KWM88, KA90, Kle91, Kli59, Kli76, KH85, Kno90, KK85, Koc57, Koc98, Koc77, Koh76, Kor64, Kor89, Kos60, Kra89, Kra57, Kre77, Kri91, Kri92a, Kri95, Kri87, Kri98, Kue61, Kuh66, Kün96, Kyl97, La 66, LaF90b, LaF96, La 78, Lac90, Lam85, Lap56c, Lap60a, Lar87, Lat91, Lau60, Lau76, Lau64, Law90, Le 77, LSB74, Led66b, Led85, Lee60, Lef76, Lei78, Lei89a, Lei89b, Lei90, LH91, Lem88, Len88, Len76, Len60, Lev77a, LH88, Lev60, Lev77b, Lev77c, Lev70]. Letter [Lev76, Lev96, Lev64, Lew61a, Lew76, Lew80a, Lew80b, Lib57, Lic77, Lie62a, Lin55, LB78b, Lit78, LPT^{+}54, Loc89b, Loe85, Loe86, LC88, Loe88, Loe89, Loe95, Lon85a, Lon52b, LP85b, Lor74, LA98, Lor57, Lov74a, Lov78a, Lov78b, Lov86, Lov89, Low76, Low85, Lyf61, Lym97, LM61a, Lyt85, Mac54a, Mac86a, Mac86b, Mac89, Mac91, Mac92, Mac98, Mac77c, Mag85, Mal65, Man66a, Man57, MM77, Mar89b, Mar93b, Mar54c, Mar54b, Mar58c, MFC62, Mar97b, Mar78b, Mar48e, Mar53c, Mar78c, Mar90, Ma’52, Mas76, Mas57, Mau77, Maw76, Maw77, Maw78a, Maw78b, May90, McA76, McA65, McC85, McC54b, McC54c, McC57, McI69, McK89, MS85, McK86, McM58, McN76, MW77, Mei51a, Mei61, Mei98, Mer54a, Mer54b, Mes97, Mey86, Mia98, Mie76, Mil62, MB86, Mil63a, Mil89b]. Letter [Mil91, Mil63b, Mil95, Min86, Min88, Mit58, Mol89, Mol96, Møl90, Mon53a, Mor57a, Mor62a, MD66, Mor66a, Mor52, Mor87b, MT96, Mor61b, Mos65, Mos85, Mot85, Mot64, Mue77, Muk93, Mül89, Mul77, Mun65, Mus78, Mus02, Mus76, Mus79, Mye90, Mye91, MJD92, NR85, NK86a, Nat74, Nat86b, Nea61, Ner76, Nin88, Nor56, Nor94b, NE76, Nov76, Nov77a, Olk77, Ols62c, Olt52, Opp49a, Opp62b, Opp90a, Opp90b, Ora65, OH85, Ore62a, Ore62b, Orm57, Orn85, Owe87, OB78, Pac59, Paj52, Pal76, Pal77b, Pan86a, Pan98a, Par88a, Par66c, Par78a, Pas51, Pas66, Pas77, Pas76, Pat87, Pat90b, Pau88, PR62, Pea90, PP63, Pen77a, Pen77b, PWW63, Pen98, Per88a, Per87a, Per92b, Pes87, Pet78, Pet92, Pfe52, PB92, Phi93, Phi94, Pip55, Pis64, Pit52a, PR85]. Letter [Pol88, Pol57a, Pol62a, Pol62b, Pon65, PW76b, Por74, PGL91, Pot49, Pow90b, Pow90c, Pow91a, Pow74, Pri58b, Pri76a, Pri86a, Pri86b, Pri89, Pri76b, Pri77, Pru85, Pya74, Rab56f, Rab66b, Rad97, Ram86a, Ram97, RL96, Rao77, Rat62, Rau96, Ray89, Ray90, Rea76, RMK65, Red74, Rei91, Rei93, Res76, RC90, Ric77, Rin61, Rin66, Rob49, Rob77, Rob78a, Rob88, Rob91a, Rob90a, Rob90b, Roc97a, RDS85, Roc97b, Rom77, Ros62, RD76, Ros98, Rot94a, Rot97a, Rot98h, Rou75, RD77, RC87, RC88, Rum54, Rus50, RGR63, RB65, Rus92a, Rus92b, Rya89, Sab76, Sac57, Sai98, Sak98, Sal78b, Sal79b, Sal75b, Sal77, San76a, Sap75b, Sau78, Say93, Say92, Sch91, Sch66, Sch92a, Sch87, Sch54b, SvH77, SB77, Sch57a, Sch04b, Sch63b, Sch76b, SN89]. Letter [Sch63d, Sch76c, Sch88a, Sco66, Sco89, Sco90, Sco92, Sco94a, Sco94b, Sco95, Sea56, Sea85b, Sel59, Sen65b, Sha65, Sha85a, Sha66, Sha92a, Sha85b, She74a, SF77, She77, She64b, She65, She85c, Shi76, Shi88, Shi66, SdSP62, Sho77, Sho76, Sho92a, SC97, Sie52, Sil65, SC76, SIF66, Sin91, Sin63b, Sip86, Sla76, Sma76a, Sma76b, Sma77, Sme76, Smi74d, Smi76, Smi85b, Smi85a, Smi87b, Smi90c, Smi94a, Smi96, Smo89, Smy64, Sob91a, Sob91b, Soc55, SD56, Sol66, Sol76, Sol77, Sol90, Spe54, Spi66c, Spi67, Spi63, Squ75, Squ77, Sta60a, Sta61, SJ64, Sta96, Sta93, Sta97, Sta87, Ste85a, Ste77a, SK85, Ste86b, Ste89f, Ste76b, Ste76c, Ste87, Ste78, Sti62, Sti52, Sti57, Sti78, Sto90, Sto97, Sto76a, Sto78b, Sto62, Sto76b, Sto77, Sto60, Str90]. Letter [SH61, Str86, SD87, Str78, Str88, Sut91, Sut77, Sut89, Swa86, Swe87a, Swe89b, Swo86, SG59, SG62a, SG62b, SG65, Szi47d, Szi60b, Szi61a, Tam71b, Tau85, Tay52, Tay56, Tay62, Tay69, Tay76, TK66, Tel54, Ten69, Tha78, Tho75, Tho87, Tho77, TW66, Tho86b, Tho89, Til64, Tir85, Tob77, Tod55, Tor77a, Tor77b, Tor77c, Tot70, Tov89, Toy95, Toy96, Tri77, Tsc87, Tsi89b, TKG94, Tuc62, Tuc64, Tuf54, Tur76, Tur87, Tyl89, Ump75, Ung78b, UC62, Ure56, Ure61, Van90a, VFIS48, Ver89, Vis59, Vog87, Voh71b, Voh71c, Voh71d, Vos88, Voy76, Wal77, Wal89b, Wal55, Wal46a, Wal76, Wal86, War95, War59, War87, Was63, Wea76, WF77, Wee95, WW61, Wei77b, WB85a, Wei94a, Wei86c, Wei46, Wei93, Wel77, Wes58, Wes76, Wes91, Wes79b]. Letter [WB85b, Whe90, Whi76a, Whi77, Whi76b, Whi78a, Whi78b, Whi53, WLP97, Wie47a, Wie78, Wie87, Wig88, Wig58, Wil95, Wil93, Wil86a, Wil89, Wil76a, Wil62, Wil75e, Wil78b, Wil86b, WA93, Win87, Wis55, Wit95a, Wit52a, Wit52b, WvH86, Wol86, Wol91b, Wol88, Wol75, Woo78, Woo85, Woo86, Woo62, Wor58, Wou76, Wou78, Yoa78, You64b, You69, You76b, You76c, You77a, You77b, You78b, You85b, You85a, You87a, You92, You87b, You77c, You78c, YC98, Zan91, Zar76, Zar64, Zei85, Zei60, Zel70, Zel76b, Zie77, Zif53, Zim87, Zim76, Zim77, Zlo47, aI98, deF51, dJ64, dW90a, vHC76, vHWD76, vL85, vL86, Ad 90, Ano62y, Ano86v, BBG^{+}85, Gaf86, Mil83b, Phy80, Sak82, Uni54]. Letters [Alb75, ARA94a, Ano46a, Ano99u, Ano99v, Ano99w, Ano99x, Ano99y, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano01k, Ano01l, Ano01m, Ano01n, Ano01o, Ano01p, Ano02l, Ano02m, Ano02n, Ano02o, Ano02p, Ano02q, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano04i, Ano04j, Ano04k, Ano04l, Ano04m, Ano04n, Ano05j, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, AG91, BED93, BPG89, BS97, BB87, BHM91, BJH68, BM93, CSF95, Cas74, Coh75a, Coh75b, CBSF95, Coo75b, CMM95, Cut75, DE95, Day75l, Day75m, DMJ98, DeN75, DRB92, DH98, DPL75, Dre46, DvH91, Dun75, Dun74, ER90, Elm65, Eps75a, FL62, For80, For75]. Letters [FK89, Fru75, GM87, Ger75, Gir75, Gla75c, Gla75d, Gol75, GH94, GM90, Gul75, HvH75, HFMH89, HFO63, HD92, HPR91, HS96, HS97, Hot75, Hur75, JG96, KS94, Kat75, KB89, Kor75, KA91, Kuh75, Lal75, LM75, LS65, LKWR89, LRD^{+}91, LC94, MD75, MC75, Mat75, Mav75, McL65, MHM90, Mul74a, Mur75, MW75, Mus75, Mye75, Ols74b, PWS^{+}96, PT87, PG91, Pri75a, PE89, Pri75d, RD75, Ris75, Rob75a, RGK95, Sal75a, Sau74, Sau75, SH89a, SH95, SB89, Sch75b, Sen65a, SH94, Smi74c, Smi75d, SW80a, SLM87, SC75, Spe75, SHT^{+}87, Ste65b, Swe02, Tel75, TA94, TM79, Tyn75, WB95, Wei75, Wes74b, Whi75, Wil75a, WS95, Wil75g, Yeo75, Zel75, AH92g, Ano65e, Flo88, HPL^{+}49]. Letting [Ano59a, Ple96]. Lev [Art71, Fel73b]. Level [DSZ74, Mei90, Bey12b, Bey12a, BR15, CMR79, Coh79, Day77j, DM18, Dro24c, Eav19, Gro82, HQL84, Jac15, Kas12, McL18, MHKL18, Mor78, Nuc18, Rot81a, Sal90a, Sca11, SF94, Ste90b, Sti18, Tre18, Van20, HI72]. Levels [Bio73]. Leventhal [Ano87o]. Leverage [Kin86, NLS25]. Leviathan [Oph73]. Levich [Ano73b]. Levine [You65a, BH65, Lev65, Mos65]. Levitation [CCH72]. levity [Old81]. Levy [Zar64]. Lew [Fel79o]. Lewin [Gol96b]. Lewis [Ber79, Dow80h, For80, Fra83, Gar97a, Kat54e, Lam69, May89, Pal64b, Scl79, She83b, Wyl56, Abr73, Eml66, Fol75, Kat54b, Ols59, Wur68]. Lewontin [Jen70, Rab70b]. liabilities [Hol80a]. Liability [Sil66]. Liaison [Lei92]. Libby [Ano58h, Lib57, Mar82b, Mat56, Pre62, Wil80a]. Liberal [Gla58, Lev64, dF79g, Sim84a]. Liberalism [Dev69]. Liberalizing [Sim54e]. Liberated [Sim46a]. Liberating [SH89b]. Liberation [Fal85, Adi69]. Liberties [Ano49l, Sie80]. Liberty [RD76, Kal62b, Wil53b]. Libya [Ark96c, Mic81, NA87i]. Libyan [Nat85a]. license [Ano93-27, Lum94a]. Licensin [Jas81]. Licensing [HG71, Rus70a]. licentiousnes [Bar76a]. Lie [Hol74, Wei94c, Ano46j]. lieber} [Ano93r]. Lieberman [And71]. Lies [Ano98k, Das66, Eve88a, Les93d, Due21]. Lieu [Mec23b]. Life [Ano94j, Ber78a, Ber04b, Bro60a, Car68, CF74, Cav48, CLR06, Dan72, Dar57, Day97, Dow81e, Fox73, Gol96b, Gom70, Gre60, Har65, Har92a, Hor71, Ive16, Lan56b, Lan58, LeR49b, Led71b, Lif69, Lla70, Loc14, Lue96, Mea93, Mul64a, Nef59, Nor74, Pan07a, Rab57h, Swe93, Vof72, Wil69c, Woo97e, Alb81, Ano96l, Ano96z, BO76, Bon57, Bon58a, Bon58b, Bor89, Com81, Cas92, Cro17d, De 05, Fel92, Isa07, Sag83, SSK10, Tha63, Yev86, PF75, dF79g, FBH81, Fer57a, Hen71a, Pla57a, Abb93, Ano90-47, Has71, Rom82, Smi83, Wil77b]. Lifelines [Woo98d]. Lifestyle [Isa93f]. Lifetime [Gru70b]. Lift [HM98]. Lifted [Ano51j]. Lifting [Hev47, Tol58]. Lifton [Eli85, Keu83, Moo95e]. Light [AM74, Bet75b, Boy86b, For86, Gol09, Ano91-34, Ark95d, Isa90p, Ner80, Rot98h, Smi90a, Wil74b, Far86, Haf79]. Lightman [Ano87n]. Lights [Sch05b, Wil92d]. Like [Alv74, Ano91-33, Ano93t, Ano98m, Fly00b, Isa00b, LaB81, Sha91a, Web03, Weg96a, Ano92r, Ano93s, Ano95l, Ano97k, Dro22h, Gol96b, Leo19, Tan88, Zag95e, Dro23j]. Like-minded [LaB81]. likely [Fin92, GB11, Mer24, Ram13b]. Lilienthal [Gre67b, Ano46-47, Ano46-81, Ano47x, Ano47u, Ano47v, Ano49-28, Ano49-31, CFU47, CF46, Day76b, HL49, Mil50, PMRP50, Pis64, And65]. Liljestrand [Kat52b]. Limbo [Kim01, Mat98a]. Limit [Aik61, Ani15, Ore62b, Riq15, Thr15]. Limitation [Ano71b, Ano88e, Cou57, Fel75b, Pin57, Bin18, Kre85a]. Limitations [Fel72a, Hum72, Lap72, LvH18, Nie55, Sco70, Sco72, Sle56, SD55, Wea75, Whi71, dRMT^{+}47, GK87, GJ84, Hol76a, Tsi84b]. Limited [Aro56, Ber75c, How62, Ing56a, Pan88, Pei82, Wal89a, Mar87a, Squ13, AM21, Cro65b, Roc81]. Limiting [BBHGA57, Rod69, Sco72, Ano88p, RH17]. limitless [Pan88]. Limits [Abb12, Alb71, Aug65, Bar73b, Ber72b, CL00, Gil85, Gol96a, HW73, Har72a, Hul78, Lap56d, Myr72, O’C87, Pog54, Rot97a, Saw80, SL58, Sei71b, Wal99, Wat73b, Wol62, dF79d, Alv80, Cav13, KKS22]. limp [Say92]. limps [NA90c]. Lincoln [You66c]. Lind [Mor87c]. Linda [Ano78n, Eri95, Fin90, Kor92, Roc85]. Lindblom [Bro79b]. Lindsay [Ano47w]. Line [Cum03, Mur55b, Wal52, Zel92, All14b, Ano95k, Bro87a, Eps97, Hoe87, Hor88b, Isa93b, Isa94a, Pod16, RM98, Sig88b, Sno80, Tho81c, Zag91b, Ada87, Cap02, Wil71, Eng91]. Linear [BW77, Wei77b]. Linenthal [Rot97b]. liners [Isa88e]. Lines [Klo85, Sch75a, Sha90, YY74, Ano11d, Bol88b, GS96, Ray80, SR81a, Sig10]. lineup [Isa85c, Isa87b]. lingering [Eri91]. Lining [Ano75h, Ark01c, Sch76b, Ano91-37, NA92e]. link [Hol78, Mec19f, Rot92s]. linkage [Eva82]. Linkages [MS77b]. Linked [War91]. Linking [Bro85d, Mye87b, Sto78a]. Linus [Sha59, All60g, Ano56l, Kal60, Kal61, Dow79a]. Lion [Ano80o, Jac74b, Van60]. Lipsky [Nan73, Pla59a]. Liquid [Abr74]. Liquified [Wol78]. Lise [For80, Gol96b]. List [AM51, Ano46-51, Ano46-52, Ano47m, Ano50h, Han00b, MR09, Tuc10a, Aft94a, AH91k, Ano97m, Cah94, NA88o, Rot04h]. listen [Ano91-31]. Listening [Gro54a, Gro54b, Gro54c, Gro54d, Isa89m]. listens [Ark89b, Was97]. Litai [May89]. Literacy [Mec21a]. Literature [Ano50-44, Mad63, Rin61, Wil81b, Wyl56, Thi46, Dev69]. Lithuania [Pau05]. litigation [Baz83]. Little [Ano90-37, Ano94s, Cum01a, Hal91, HLL03, Leb91b, Nor04, Pod01, Sch75c, SG75, WS91, War89, AT88, Ano88g, Ano92q, Ano94j, Ano95n, Ark93a, Bou80, Cor95, Gar97c, Gea89, Han89, Isa89l, Isa90b, Mil89a, Moo95j, Moo96m, PC92b, Rot95b, Smi90a, Tar93, BT94, Joh97c, Mol96, Sei84]. Livable [Pri74, Fel81e, Tho81b]. Live [Dow82e, Har72a, Mac54b, Rab60g, Szi60a, Szi63b, Bro79d, Bro79e, Mat80b, Smi80]. lively [Ber87a, Les93b, Wil77b]. Livermore [Yor75a, Bat88, Blu88, Gla78, Zam95b]. Lives [Ber86b, Gen84a, Gin76, Lan58, Whi78b, Wil71, Bea86b, HBD17, Fer57a, Nei82a, Pow90a]. Living [Can72, Dob54, Ger72, Har76a, Jon65, Mos92, Rab55c, Rab72g, Ano96l, Jr22c, Dro22h, McK23b, Tho80d, Ded60, Fin90, For83, Fre84b]. Livingston [All55l, Dow83c, Wil74b]. Lloyd [Car72, Sai72]. LMFBR [Jac74b]. LNG [Hav04, Sch79a]. Loader [Dow83e]. Lobby [Har92b, Ano87-40]. Lobbying [Hal71, Pai85e, Rot97c, Smi74b, Ano86f, Har96]. lobbyist [Isa83c]. Local [Ano48a, Deu68, Gor51a, Moo51, Zei51, Kib12, MHKL18]. Locall [Lap59f]. Location [Ano92u, Aug48c, MEA^{+}47]. Locations [NA95a]. Loch [NA91l]. Lockdown [SS06]. Locked [HVK^{+}06]. Lockheed [Kal78]. Lodestar [Rot97g]. Loeb [Lon82, Swe88]. Lofty [Boy00]. Logic [Arm68, Mar02c, CBSF95, Fal85, Pif24, Pol79b, SF94, Wes79a]. Logistics [dQR70]. Logos [Mar64a]. logs [Ram82]. Logue [McK74]. London [Ano46-53, Ano55y, Ano80o, Ano81i, Ano98p, Aur71b, Aur71c, Aur71a, Aur71d, Aur71e, Aur71f, Aur71g, Aur72b, Aur72d, Aur72c, Aur72a, Aur72e, Aur72f, Cal60a, Cal60b, Gol73a, Gol73b, Hes55, Joh81b, Mos80, Mos89, Pat78a, Pat78b, Pat79c, Pat79d, Pat79e, Pat79f, Pat80a, Pat80b, Pat80c, Pat80d, Pat80e, Pat81b, Pat81c, Pri70, Spi57a, You63b, Gri83a]. Lonely [You77c]. Long [Abb04b, Ano86i, Ant74, Ark87f, Cra70b, Dow81a, DL21, E^{+}84, Fel57, Hir89a, Hol73, Kri85, Lov74c, Mei51c, Rab50a, Sim56, Smi73, Szi60a, Tsi75b, Wis55, WvH86, AT88, BS97, Col95, DM18, Due21, Eng91, Gla12, Gom82a, GH83, Ial18, Mal79, Mes97, Nil14, RKK22, RM98, Sil79b, Twi23, vHC86, AH02, Alv01a, Bos06a, Fly01b, Pra06c, Mei51a]. Long-duration [DL21, Twi23]. Long-Range [Hol73, Rab50a, Smi73, Szi60a, Tsi75b, Wis55]. Long-Term [Ant74, E^{+}84, Kri85, Lov74c, Mei51c, WvH86, DM18, Due21, Ial18, Mes97, vHC86, Mei51a]. Longage [Har88b]. Longer [WW60, Lan79, Nit80, Rot77]. Longworth [Nad98a]. Lonsdale [All54b]. Look [Ack90g, Ali57, Ano95k, Arn98b, Cra70b, Ded65, Ell93, Fel57, Fre72a, Hal75a, Kuy54, Lea69, McC69, MM71, Pri58a, Shr23, Sim56, SNB^{+}54, Str05c, Way65, Ano83p, CN91, GL90, Har10, Has88, Hui87, Isa82d, MM11, Nat88, Swe85, Toy84b, WL84, Zag95e, vHJ96, Pie16, SNB^{+}54, Ano46t]. Looking [Ano89t, Ber01, Cum99a, Dro17, Fel72b, Mei64, Puz93a, Rot96h, SM82, Sch69b, Wei85c, vH77b, Pai83a, Smi91, Squ17b, NA93j]. looking-glass [Squ17b]. Looks [Bur74, Kna54, She64a, Ten75, Zei50b, Zei51, Ano79c, Cha87, Lak78, Mec15b, Sap80a]. loon [Ano97j]. loop [Ano94o]. Loophole [Bur90b]. loopholes [Ano96l, Cha87, Moo97p, Pai83a]. Loose [Bla20, Cho01, Ber89a]. Lopez [Ano97p]. lopsided [Sha91c]. Lord [Ing56c, Ano51-29, Mar58c]. Lords [Che52, Shi48b, Sim59a, LLR^{+}47]. Loren [Soy89]. Lorenz [Rab66b, UM67]. Lorna [You76a]. Lorrain [Dow83b]. Lose [Bat52, Fra80, Rab52b, Ano90-27, Dro19, FP17, Isa89k, Kib12, Nay12, NA90k, Sap80b, Squ12, Zam93]. Loser [La 91, Rot93p]. Losers [Sch69c]. loses [Isa86d, Rot94d]. Losing [Pau05, Sch04a, Ros98, Sie24, Bar90b]. Loss [Lew70b, Pod01, Dro22f, Isa88e, Moo25]. Lost [Ack99, Ano89u, Ano98l, CG12, Cir05a, Cof67, Con69, Dav74, Day75b, Gom71, Gou46, Isa05d, Jan76a, Joh80b, Kra94, Kul06, Lig84, MD75, Pow91a, Rei72, Sau75, Wol75, Mak98, Mos89, Sut95, Alc81, MJ81, RGA81, Ano86l]. lot [Dro24e, Fie19, Zag95e]. Lots [Sch00a]. loud [Gra85]. louder [Ano12a, Dow84d]. loudly [Fei90]. Louis [Ano84d, Din69, Don86a, Fel82h, For80, Gan84, Her90b, Nei82a, Sha59, dF79f, Ano46-70, Kal82h]. Louise [Bal79c]. Louisiana [NA92a]. lovable [Ark93a]. Love [Boy05, Dow83b, Rus52a, Sti67, Wes03, Isa89k, Isa91r, Mos89, Rot83, ZA04]. Loved [Con69, Mea72]. Lovejoy [Dow77]. Lovell [Coh69]. Lovely [Wat92]. lovers [Ano93x]. Lovins [Nel84, Nef78, Nel84, Dow83c]. Low [Bio73, Bar90a, Bas80, Boy00, HQL84, LC09, Llu99, Mei90, Mye75, Sal90a, Ste90b, Wil92a, Ano97m, Bey12b, Bey12a, CMR79, Coh79, DM18, FR97b, Gro82, Gun20, Hil12, Kap92, Kas12, Mat98b, Mat12, Moo15, Mor78, NA89a, Pai85c, Rot81a, SF94, Sin83, Tho12, Zik21, PS04, HI72]. low-carbon [Gun20]. Low-Cost [LC09]. Low-dose [Wil92a, Hil12, Tho12]. low-enriched [Moo15]. Low-intensity [Bar90a]. Low-Level [Mei90, HQL84, Sal90a, Ste90b, Bey12b, Bey12a, Coh79, DM18, Gro82, Kas12, Rot81a, SF94, HI72]. low-yield [Zik21]. Lowdon [Fri64]. Löwenstein [Wri63a]. Lowenthall [Dow81i]. lower [Ano11h]. Lowered [LaF90b]. Lowering [Moo02a]. Loyal [Rot03d]. Loyalty [Ano50w, Ano50x, Ano50y, Ano53-40, Ano55j, Ano55m, Bro55c, Cle48a, GL48, Gro51, MB48, Mor55, Mur55a, Nat56, O’B48, Rab51e, Shi55, Spi49, Sta48, Ste53, Whi51b, Ano54w, Ram13a, Ano47h, Rus51a, Gre58]. LSD [Lie62a]. Lt [Lap55a]. Lt. [Sch06c]. Ltd. [Woh78]. Lttle [Leb91b]. Lu [Ols62a]. Luard [Alc63]. Lubrano [Ano78n]. Lucas [Pat81a, Som91]. Luchkov [Leb91b]. Lucid [Ago77]. Lucky [CO03, Mar58a, Ano94n, Sal78a]. Luddite [Tsi78b]. Lueders [Day97]. Lugar [Loc93b, Loc95, Lug00]. Lull [Ano46-37, Mar65a]. Lulu [NA89i]. Luna [Lew69b]. Lunar [Ano69o, Foz62c, Har69b, Kui73, Lat73, Lew69a, Mic69, Mid73, O’K69, Sch73, Szi71, Tur73, Vol74, BB22, Smi72c]. lust [Rom92]. Luther [Law88]. Luttwak [Mil86b]. Lydia [Ano86i, Dow84d]. Lying [Ber70b, Moo94g, Nor94b, The89c, Zag95c]. Lyman [Par83, Phi84a]. Lynn [Pat81a, Mil86b]. Lyons [Ano71g]. Lysenko [Cra49, Dob52, Rab59h, Hec70, Rab72a]. Lyttleton [Smi58f].

M [Abb85, Abb93, Ada52, Ada61, All55l, Ano60f, Ano62l, Ano78m, Ano79a, Ano85c, Ano86j, Ano88x, Ano14a, Bei84, Bri86, Bro57a, Bro57b, Bro54, Bur90c, Cal84a, Cav57, Fal61b, Fla90, Gol69b, Hal93a, Jac80b, Kat52a, Kat52b, Ken71, Kin86, Kre90, Lal67b, Mei90, Men92b, Mil86b, Moo96b, Moo96c, Moo99g, Mor82a, Mor85a, Mot68, Pan63, Pap88, PF75, Pow89b, RK89a, S.80, Sch63c, Sco74, Sha81a, She79, Sid69, Sta85, SB91, The89e, Ull63, Van03, War92a, Wil76b, Wil77b, Wol91b, Wri63a, Wyl56, You66b, dF79h, dF79e, dF79g, Cra67a, Har90c]. M-16 [Har90c]. M. [Eme54]. M.I.T. [Ano69a, Sal69]. Mabel [Fal61b]. Macaroni [Kir90]. MacArthur [Ano52h, Cum92a, Mac51]. Macauley [Bar67a]. MacFarlane [dF79g, Ano12g]. Machine [Coo74, Fle70, Gus95, Lan83a, Mar73, Mos67, PPE22, Ano94l, Dro17, NA87j, vH89, Dow81d, Liv69, Sch71a]. Machinery [Kal87, Min55, Sch86]. Machines [Ada67, ARA94a, All16, Tel47b, Bel93]. machismo [Isa89l]. Macho [Dav77a, Ack91b]. MacKenzie [Pit91]. Macromolecules [Eis72]. Macy [Mes83b]. MAD [Buc81, BCB81, Chi82, Cou81, Fel80f, Gue79, Jer85, Lin82]. Madame [Alb54, Ble54, Hai54, Mer54a, Rab54c, Rum54, Tod55]. Madder [Fel80f]. Maddox [Fre76a]. Made [Ano78m, Ark00d, Lan61a, Mel58, Van60, Vog90b, Ano93-27, Ano96j, Bra95, LP10, Moo95d, Nov77b, Rot95l, Tin23]. Madison [Jac71b, Isa90m]. Madness [Ark02a, Ing68a, WC73, Cam95b, Lag80, Moo95g]. Madonna [Ano92v]. Madrid [Ano87n, Ano89e, Ano91u]. mafia [Zag95a]. Magazine [Bre93, Ano91-52]. Magic [Cal61b, Fre57, Ner69, Pol56, Pol57a, Sol66]. Magical [MZM01]. Maginot [Ada87, Sno80, Wal52]. magistrature [Bug78]. Magnetic [CCH72, Par73]. Magnetism [Fre72a]. Magnets [All59l]. Magnitude [McV71]. Magnusson [Bla84]. Mahathir [Ahm04]. maiden [Isa93g]. Mail [Wil62, Fer03]. Mailer [Rou75]. Main [Rab50d, Wil76a, Car03]. Maine [Woo01a]. Mainholf [Wil81a]. Mainland [Esp72b, Lal64c, Mur72, Rot72]. Mainstreaming [La 88a]. Maintenance [NA91h]. Majesty [Moo01f, Opp03a]. Major [Ano49-29, Ano54q, Ano87-30, Boy89b, FKM87, Har87c, Iri79, Lap69, Sak84, Shu78, Smi83, Tur51, Tur52, Ble88, Cim84, Cro18, Mec20h, Wai60]. Majority [GM54, SZC54]. Make [Ano93-27, Dys75, FO07, HvH75, Hul76, Moo95i, RR07, Say84, SL87, Dun85, Eav18, GB77, Gra89, Hir89a, Isa89k, Kra92, Mec21f, Nei82b, Pan92, Sha95c, Squ12, Sto23]. Maker [Bro48, Fed76]. Makers [Zhe90c, Dro17, Kis76a, Ahe91, Ano81f]. Makes [Isa90t, Isa91a, Sig98, AGB22, Ial18, Isa88f, McK22, McK25a]. Makhijani [Pre96]. Making [Ack88n, AP88, AT88, AH91h, AH91i, AH91g, AH91j, AH91k, AH91f, Alv01b, Ano46s, Ano54i, Ano54k, Ano96o, Buk02, CAN88, CNB95, Cra84c, De 74, Dow81a, Dow83a, Fel75c, Gal13, Gol95a, Han95a, Hor88b, Jan81b, Kow55, LL12, Mar06a, Mes87, Moo94a, MJ80, Pai02, Ram88a, RS88c, Rho95, Ros08, Rot93m, Sal71, Sch54a, Smi68, Sob21, Ste88, Tod79, Tsi75a, Wag88, Whi71, Wil93, Yor88, Alb87c, Dal83, Hol76a, Hol87b, Hou85, Isa95e, NA88o, Ros78b, Rot98d, Scl82, Wei18, Yem87, Ano78n, Gib82, Pat81a, Mes64b, Sar85, Tay87]. Malaysia [RR81a]. Malaysian [Ibr13]. Malcolm [Han98a]. Male [Ano91-44]. Malevolence [DeN74b]. Malevolent [Coh75b, Sch84b]. Malfunctioning [Wil75b]. Malicious [Yun94]. maligned [Hig89]. malpractice [McC93]. Malthusian [Ing71, WI71]. Man [Ano61n, Ano69m, Ano70e, Ano98k, Bac58, Bea64, Ber58, Bor57a, Bor63, Bro55a, Bru58, Cha58a, Cle71, Coo77, Cro58, Dan72, Dix65, Dje72a, Ein51, Fed72, Foz63a, Fra61b, Fre72b, Fri78, Gee66, Gen84c, Gla56b, Gol69a, Gol69c, Gou63, Guh65, Har69a, Hau64, Hoa72, Hol58, Hor76, Hul76, Hux59b, Hym69a, Kat52b, Kne65, Kuz58, Las62, LS58, Lea07, Leb91b, Lib58, Lib65, Lil49c, Liv68, Lop68, Lov69, Mid65, Mit54, Mul68, Nac72, Ner76, Neu58a, Neu58b, Oga58, Osb58, Pie59, Pla65, Rab56k, Rab56l, Rab56j, Rab61g, Rab62f, Rab63f, Rab69d, RL69, Rab72b, Ram72, Rev66, Rob91b, Rus56b, SL58, SE58, Sem59, Sto59a, SG69b, Tam74, Tel75, Ter58, Tod81, Wag61, Wat63, Wat73a, Wel60, Wes55, Wid61a, Wid61b, Wol65, Woo55, Wri63a, Wyl53]. Man [You58, dF79f, dF79h, Ano91-27, Ano94l, Ano97k, Car03, Dro23l, EB78, Fre93a, Har92a, Hil16, Hil20, Hoa77, Kis93a, LS94, Lan98, Law75, Nor02b, Nor03, Pra93a, Rot93h, Spe95a, Ste77c, Tin23, Wur92, Har50b, Hoa72, Joh03, Cro67, Gal54, Has71, Mei77a, Nor04, Rei72, Sim62, Nor64, Sil71, Woo65]. man-break [Fre93a]. man-made [Tin23]. man-man [Rot93h]. manage [Gei16a, MvE13, Pan22]. Management [Ano78n, Ano82f, CR71, Das96, DSZ74, Fel83d, Fox61, HG84, Hin69a, Kne65, McE70, Men68, She64c, Zin79, Arc86, Ble87, Joh83b, Kah86, KM18, SJ82, VU22, Wil85a, Bla84, Fuc70]. manager [Wil79]. managers [Dro21b]. Managing [Akh17, Ban11, Car17, HV18, Kre93, Lon85a, Sar17, Ano78n, For87]. Managua [Win90]. Manchester [Bri80b]. Mancuso [AR80]. mandate [BOW10]. Mandatory [Gol97, Moo14]. Mandela [Mas99b]. Mandelbaum [Mar78a, Roc81, Gan83a]. Maneuvered [Sch00c]. Maneuvers [Bar80f]. maneuvres [Ada86c]. Manfred [dF79d]. Manhattan [Alv87, Gol88b, Moo95d, Ano46n, Ano46-51, Ano46-52, Ano47m, Dro23b, Ell92, FS95, Hut46b, Mye91, Nor02b, Nor03, Rho07, Rho15, SB23, Smi68]. Mania [Fly02c, Flo81, Nol90, NA87h, Pai82b]. Manifesto [Rab68c, Rot81b]. Manipulators [Sch69c]. Mankind [Ale69, Ano46-53, Gre63b, Hel77b, Jon63b, Kel78, New58, Rus53, Rus54, Kel20, Han75, Tor61]. Manmade [Scu65]. Mann [Ano10c, Bar64a, Mar54e]. Manna [WC73, Ano93t]. Manned [O’K69, Lov66, Smy64]. manners [Lan89g]. Manoevre [Dow80h]. Manpower [Ano50z, Ano50-36, Ano51a, Ano51s, Axe51a, Kat51b, Kat52c, Kat52d, Kat53b, Mey54, Sac48, Smy51, Ano56c, Kat57b, Kat57b]. Manta [Fly05a]. Manual [Wol55]. Manuals [Ano50-41]. Manuel [Bel93]. Manufacture [All57k]. Manuscripts [For80]. Many [Arn92b, Gor11, Gor20, Hub57, Kah59, Lap76, Rol57, Sau01, Sim59e, Ano91-37, Ano93u, Ark97a, Bea86b, Dun85, Fie24, Kre93, PC92b, Ros88, Wei76b, vH16]. Mao [Cum04, Jen74, LM76, Old69, Sim88]. Maoism [Gra69]. Map [Rab57f, TW66, Tsi75a]. mapmaking [Yun94]. MAPS [Mey79]. March [Ano58g, Lei47, Rab54a, Wit03a, Pow98, Sil79b, Soa96, Ano47f, Ano47v, Ark02a, Bor53, Fel93, Sal69, Tuf54, WG69, Wit03a]. Marco [Ols62a]. Marcus [Mul65]. Margaret [Bal79c, Dow80i, Pla64a, Sut54, You76a]. margin [Pai82d, Rau96]. Marie [Nor64, Rea80]. Marilyn [La 91]. Marine [Che70, Pat75, Mer24]. Mariner [Lew65a]. Mariners [Cra72]. marines [Ano90t]. Maritime [Dra06, NA88f, Ano87i, Bar87, Pos87]. Marjorie [Bar67b]. Mark [Bar90c, Cal82, Dow82g, Jas81, RK89a, Ses83, Sti83, Wei89a, Ack90c, Rev92, WSG^{+}15, Dro24d]. Market [War68, Bar81c, Bel94, Ber78b, BMPH18, Cul14, Hus90, Kim22, Kla88, Law11, Pan23, Zag94f]. market-oriented [Pan23]. Marketing [Ahm00c, Rot00a]. Markets [Bro79b, Com08, Net71, Ano89t, Har87d]. Markey [Gaf86]. Markovits [Ano81c]. Markus [Pat82a]. Markusen [Gol92a, Gol93a]. Maroncelli [Van03]. Mars [Foz63a, Hor71, Jac71a, Lew65c, Lew70c, Sag74]. Marschak [Lof51]. Marseille [Mar61c, Rus58a]. Marsh [Gre82b]. Marshall [Mil88, Tho80g, Wil80a, LP92, Mas48a, Prz82, Sto19d, Str56, Woo98c, Woo00d, Woo00e]. Marshallese [Bro70]. Martin [Bri86, Bro20, Cah76, Cra88, Dow81g, Fly98a, Gen86, Gor71, Hyd75, Nei82a, Roc81, Rot97c, Sim65, Sta64, Ste89e]. Martyn [Rei78]. Martyrs [Pol91b]. Maruki [Mes83a]. MaRV [Dow80e]. Marvin [Kar80, Sch68f, Wal84]. Marx [Rab73d, Ano94-28]. Marxism [Cra83b, Fre57, Pol56, Pol57a]. Mary [Ano86i, Dow80e, Dow82c, Hol83a, Hur78a, Kre83a, Moo00a, Per91a]. Masks [Mar89e, Dub85]. Mason [Ano78n, DeN74a, Fre76b]. Mass [Ano74d, Ano78n, Bro60b, Com51, Hsu79, Kra10, Pet54, Smi61, Smi05, Yor74, Zil64, Dor84, Rob78c]. Mass. [Ano80b]. Massachusetts [Nor02a]. Massive [AK95b]. Master [Sch07b, Sta58, Wei89b, Wil74b]. Masterminds [Lew65c]. Masters [Alb98d, Loo50, And55, Mag81]. match [Ano96j]. Material [AK04, Don88, GM09a, Pin57, AP88, AH91g, HD23, Rot77, Wag88]. Materialism [Bix49, Bri49, Hol53, Mar49b, Mot64, Sta49]. Materials [ABB^{+}46, Ano45c, Cog82, Sch74a, Sch62b, Str06b, rWeKHF47, Wil75b, DM24a, Epp20, GM14, Moo14, Rot97e, Sie80]. Maternal [Tic94]. Math [Wil76b]. Mathematical [Hyd75, Ney73, Dav69]. Mathematician [Wie56, Wil76b]. Mathematicians [Ano65g]. Mathematics [Kli52, Mor59c, Joh58, Pal84]. Matheny [Eav17b]. matrix [Mec18c]. Matter [Alb00, Ano74e, Coh93a, Dar71, Dob54, DuB48, Fel61, GNMS54, Key57, Rus51b, Ste76c, Str06f, Tel73, Wei68a, Wie78, CH54b, Day79, Man84b, Rev21, Rot92n, Rot95c, RC88]. Matters [Dal03, Kat80a, Mar65f, Pai02, Moo91f, Mor87b, Smi16]. Mattfeld [Sin67a]. Matthew [Rez07c]. mattress [Alb94g]. Maurice [Jac73f]. Max [Ano56b, For80, Smi61, Bor65a, Bor65b, Bor65c, Bor68, Tsi84a, Wil67a]. Maximalist [Bor48]. maximize [Coo78]. Maximum [All55b, Ing85]. Maxwell [Har60a, Pow90a]. May [Ano55y, Ano84k, Com74, Dje74, Kou75, Mel75, Rob08, Sch47, You69, Ano94s, Ano96m, Ano14g, Ark83a, BHK^{+}21, Cla77, Coa47, Dah77, DS47b, Fie24, Flo88, Har89a, KGS25, NA90k, Pra22, Ros94, She85b, Ste77c, Alb98c, Ano45k, Ano52h, Ano78q, Ano80n, GM54, vMC98, Las62]. Maya [Dro22a]. Mayak [Lar99]. Maybe [Ano94n, Isa98e, Ano91-37, Ano92s, Fly98c, Isa95d, Tiu94]. Mayberry [Dow82b]. Mayday [LaF86]. Maydo [Fra61a]. Mayer [Dow83c]. mayhem [Mos91]. Mayor [Zei50b, Eav19, Zei53]. Mazarr [Moo98b]. Mazmanian [Yar80a]. McCarran [Ano52m, Ano87-27, Cun52]. McCarthy [Dow80a]. McCarthyism [SB77]. McCartney [Mos92]. McCastro [Ano91-39]. MccGwire [Bla87a]. McClintock [Tod79]. McCloy [Ano78n, Ano47n, PT87, Ras83]. McCone [All58a]. McCormack [Abr76]. McCormick [Sch90]. McDougall [Hol86b]. McFaul [Mec16c]. McGovern [HM72]. McHale [Fri70b]. McKean [Bou61]. McKellar [Ano47x]. McKibben [Ano12a, McK23a]. McKinnell [Tho80g]. McKinney [Ano60l, Ham61c, Kaz61]. McKitrick [Mil85c]. McKnight [McG84]. McMahon [RDS^{+}52, Ano46-55, Ano46-54, Ano46-56, Ano46-57, Ano46-85, Ano46-91, Ano46a, Ano47w, Coh48, DeC48, Fox52, Opp49a, Rab52g, Tru46, Wal46b]. McMoney [Ano91-40]. McMurin [Saw81b]. McNamara [Got87, Kap93, Mar65b, MF63, Mea93, Sha95a, Sto66b, Kap67, Mea93]. McNaugher [Ste86c]. McNeil [Kih47f]. McNeill [Sto55a]. McPhee [Fol74, G.84a]. McT [Wur68]. Me [Hir02, Str06h, Ano98p, AH89a, Cas77f, Leo19, Moo93k, Ark98d, Say84]. mea [Ano96o]. Mead [Ano78n, Bal79c, Sut54, Dow80i, Pla64a]. Meadows [Cha74, Gol73a, Cha74]. Meager [Phi94, AK95b]. Mean [Kol64, Ten69, Eva75, IS92, Isa88e, Isa94c, Kri91, McL18, Ste86a]. Meaning [Har60b, Moo00d, Mul60, Tho80g, Wei78d]. meaningful [Squ12]. Meaningless [Bet81a, HE76]. Means [Kin57, Mah59b, Oli48, Wig61, Yor74, Fel81h, Ger13, Hec12, Hec20, Ibr13, Lev12, Loc91a, Nee79, Nel92a, Ric98b, Rot93f, Wes81, Zer13]. meant [She85f, Wei83c, Dow84c]. Mearsheimer [Ano85b, Alf84]. Measure [Ano63s, Aug48b]. measurement [Ano82q]. Measurements [Ano62r, RSW55]. Measures [Ano55j, Bro55c, Cav53, Mor55, Par53, Sal90b, HK21, Len11, dTdLF24]. Mechanics [Adl69]. mechanisms [RR81b, Wol82, Sha81a]. Medawar [Gil85, Key81]. Medford [PF75, Shi54c]. Media [Goo84, Hal93a, Ame86, Ano96m, Dor85a, DH86, Fri11, Man84c, Sox80, Tis90]. Mediation [AB77]. Mediative [Zim75]. Medical [Abr80, Abr84, Ano46-58, Ano56h, Ano91-45, Bar48, Bro74, Cha58a, Cog82, Dow80d, Dow81h, Dow83b, Fre72a, GB11, LS58, Lep62, Par62, Rho46, Uni57, War48, Wil66, Fin77b, Hag77, Nat77, She14, Osb58, Hag77, Has57]. medically [Hau17]. Medicine [Ano50a, CM84, Cla56, Esp72a, Ist98, Lea70, Tod81, LeR50a, Sil79a].

Mediterranean [Bar98, NA94f]. medium [NA87j]. Medusa [De 70b]. Medvedev [Ano78o, Rab72a, Ano78e, Hec70, Mar91a, Par80]. Meehan [Hun85a]. meek [Pet96]. Meet [Ano46-49, Fri56, Hir55, Cra84g, KK12, Lan90d, LBCA23, Moo94b, Bro47c]. Meeting [Ano51i, Ano51y, Ano84k, Atl47, DJ13, Kan69, Pom65, Rub74, Sko68, Ste01, Eps91, Fra87, Lie46, Soc55]. Meetings [Ano46-48, Boo46, Bor46, BH53, Foz61b, Rab51c]. Meets [Ano46-35, Ano46-41, Bra09, Isa04b, Bul03b, Ano95k, Azh93, CW81, Har92a, Isa89i, Sto96, Zag94f]. Mega [Lan02]. Mega-Problems [Lan02]. megaton [NA89a]. Megatonnage [Ikl75, NA89r]. Megatons [Lan02, YC98]. megawatts [YC98]. Meier [Pla57b]. meinen [For80]. Meitner [Gol96b, For80]. Mel [Uhl98b]. Melman [Cav58, Mos91]. melt [Mer24]. Meltdown [Day75b, Man04, BvH82, Wal83a, Sti83, BW83]. melting [Bra17, Hus24]. Melts [Ark00a, Dro19]. Melvin [Saw81a]. Melvyn [Per94]. Member [Ano46-43]. Members [Ano46j, Ano46-81, Ano52g, Ano85w, BWS^{+}80, CG04b, Law84b, Lor70, Sau17]. membership [JC54]. Meme [YSW21]. Memo [BBD^{+}46, Ger21]. Memoir [Gen86, Gol78a, Nol17, Wei18, Kre90]. Memoirs [Ano56g, Her83a, TS01, Wil71, Hol91, Pal02]. Memorandum [Ame47, Ano63p, Ano68f, Ano71d, Bar46b, Sal70]. Memorial [Ano52-28, Ano80o, Ano94j, Bul02f]. Memoriam [Ano73d, Com48b, The48, Fel73b, Moo95l, Moo99e, Hol90, RF79, Wei86e]. Memories [Dro23b, Fri20, McM94a, Fri95, Isa90r]. Memory [Elm74, Pai04b, Shr70, Ano86z, Sco94b, Sch81a]. Men [Ano56d, Aug56b, Cha75a, Hau64, Her90a, HM72, Hus03, Lew61b, Lew68c, She89, Tho80g, Abb89, ARA94a, Ano93-32, Ano95l, Gur75, Oli89, Wur94b, Lan55, PMRP50, Ada67, Kap67, Mar69]. Menace [Ano52g, Fel74b, Har54, Ims60, Spe54, Sut77, Kre88f, Yor75b, Met83]. Menachem [Smi97]. menagerie [Fra87]. Mended [Kin67]. Mendel [Ada87]. Menem [Est01, Gol90c]. Menges [RK93]. mental [GG16, SS24, vV23]. Mentality [Ein47, Sie04b]. mention [Ano86o, Ano98p]. menu [Dea88]. Meow [Ano98l]. Mercedes [Dow84d]. Merchant [Ano55g, Veg97]. merchants [Kla77, Kla85]. Mercury [Lew63a]. Meredith [Mar98c]. merger [Mez12]. Merits [Kar88a]. Mermelstein [Sch68f]. merrier [SD15]. Merrimack [Wol65]. Merritt [Ano85c]. Merry [Ram64, Ros75]. Merry-Go-Round [Ram64, Ros75]. Merton [Ull63, dF79f]. Mesa [Tay98]. Mesarovic [Han75]. Mescalero [Car94c]. Meselson [Rez07c]. Meserve [Bul09i]. Mess [Dav94, Hin69b, Hua02, Ano95t]. Message [Ahm00c, Ano48-33, Ano58d, Ein48a, Eis54b, Joh67a, Joh67b, Ken62, Lew65a, Wei59b, Ano79d, Ano89v, Bak78, Kis81, Pan98b, Sak83]. Messages [BBH^{+}73, Ken61b, Khr61, YTS61]. Messieurs [Kas75]. Messy [Ack90h, Moo93g]. Metal [Abr74, PR14]. Metallurgy [Pfe51]. metaphor [Ack91e]. Metaphors [How65a]. Metaphysical [Mil68]. Metarchons [Wri65]. Meteorites [And72]. Meteorological [Rie60, WJ61]. Meteors [Sch60a]. meter [Ano90-31]. Meters [Tsi75a]. Methane [Dow83c]. Method [Nie55, Say55, SD55, Tur73, Wil59, Lov78b]. Methods [Ano81f, BCH^{+}48, Nit57, Bol88c, Coa47, Sch95b, Kat52a]. Metraux [Sut54]. Metric [Fra74, dF79f]. Metropolis [Mei70]. Metropolitan [Rei63]. Metz [Ano86k]. Mexican [Gue14, Tél68]. Mexicans [Rot93f]. Mexico [Ano79d, Fly00c, Gol11, Mir89, Nad04, Wio79a]. Meyer [Pla57b]. M’Gonigle [Jas81]. Miami [Tre18]. Mice [Rus56b]. Michael [Ano78n, Ano86i, Bla87a, Boy05, Can70, Chi78, Con67, Dow88, Dow79d, Dow81c, Dow83a, Dow83f, Dow84b, Dow84c, Fra83, Gan83a, Har79b, Har86e, Her85, Jas81, Jer87, LaF86, Leb91a, Lon85c, Lyn69, Mar78a, Mar59a, Moo95b, Moo98b, Nan73, Rab72a, Roc81, S.80, Sim84b, Som91, Tel76, Uhl95b, Wal83b, dF79g, vH77a, Ano09, Ano10c, Dro22i, Mec16c]. Michaelis [Ano78n]. Michel [Dro22f]. Michele [Gra93]. Michelson [Wil74b, Dud75a]. Michigan [Hor66]. Michihiko [Bru55a]. Michio [Gil83]. Michurinist [Dob52]. Mickey [Lan90d, Smi90c]. Mickiewicz [Eva86]. Microbiology [Ims60, Jac68]. Microchips [Ban22]. micromanaging [Isa89g]. Micronesia [Joh79, Woo98d]. Microscopic [Gal54]. microwave [Sol79]. Microwaves [And78a]. Midas [NA92k]. midcourse [Ano18, Gre18]. Middle [Gan03, Jon90b, Mye89, CS16, AS16, Bad66, Bar10a, Ben02, Bin67, Bin84, Bou11, Bro16, Bur62a, CAM20, Dro16b, Evr78, Fal79, Fel78a, Fel81i, Fox48, Har14b, Hay79, HGV10, IS15, Jon90b, Kla91, KJE16, Kum95, Lan14, Mem79, Pol67, RM16, Ram04, Sai14, Wil86b]. Middle-Run [Fox48]. Mideast [Gre07]. Midgetman [BR85, Cim84, Isa88f, Mag88b, WW86]. Midnight [Ano74g, Ano02j, Boa07, Bri86, Day74f, Dow82c, IF51, Jac74b, KFM50, Moo95m, Ack88h, Ack90a, Boa98, Fel80b, Fel81d, Fel84e, Hur88, Rab50e, Sas17, SM81, Ste97, Mas84, WM84, Wei55b]. Midwest [Ano46-59]. MiG [Ano91-36]. might [Ano88g, HS13, Man77]. Mighty [NA88g, Isa90k]. migrating [Hib95e]. Migration [Bil86, McL18, Hol84c, Tho80d]. MiGs [Ano94n]. Mihajlo [Han75]. Mikaelian [Dro22i]. Mike [Dow81g, Dow82f, Sti83]. Mikhail [Jac92, Rab65c, Kra91]. Miksche [She60]. Milan [Bev56, Shi56a, Top56]. Mile [Ano80k, Col79a, Fai79, Fel79l, Fri11, Gil80, HK13, KKHK79, Man04, Sti83, Wal83a, Ano13f, Wal85b]. miles [Rev81]. militaries [Kla97b]. Militarism [Siv75, Bar92a, Tsc87]. Militarization [Cra84f, Sie85, Tir85, Zit17]. militarized [Wie85]. Military [Abb85, AW85, Ano46-60, Ano48n, Ano49-30, Ano68j, Ano88d, Ano90p, Ano94m, Axe51g, Bar81f, Ber53a, Bet70, Bir87, Bon49, Cle48c, CHP71, Cae73, Cai89, Cai90, Che88, Cra83b, Din78, Dou69, Dow81a, Dum75, DW90b, Ede58, Ein47, Eps75c, Fin65, Gal56, Gol91a, Goo86, HB63, HI05b, Hen84, Hig48a, HAIT80, Hul98, Ing56a, Isa90z, Isa94g, Isa95a, Isa95b, Isa95c, Isa96e, Jac80b, Jan64, Jon87, Kyo85, Kal77a, Kar80, Kat52c, Kay49, Kno66, Lac85, Lan90e, Lan90m, Leg56, Lev64, Lyo63, Mac65a, Mar89a, Mar54b, Mar89e, McG90, McN63a, Mel71, Mor54, Mor87c, Mor62d, Mur84, Ode88, Opp51a, Rab56l, Rep85b, Ric89, Rid47, Rod70, Ros89b, Rot92k, Rub63, Sch80e, Sco71, Sea85b, Sen63, She56, She61, Sin63a, Siv80, Soh64, Som91, Spo64, Ste46, Ste86c, SD87].

Military [Tam71b, Tar93, Tay76, Tsi72b, Tsi77d, Wil78b, Wol71a, Wol71b, Wol86, Wri63a, YG77a, Abr83, AG86, Alb84, APvH86, AF88, Ano87-42, Ano91-40, Bau84, BMPd89, Brz12, CA87, Coo90, Cot22, Cum97a, DeG84, DE23, Dum86, Fal85, FP17, Gal81b, Gar88b, Gar77b, Her78b, HK02, Hoe87, HSJH18, Hol80b, Hor14, HK21, Isa80, Isa82a, Isa84b, Isa86b, Isa86e, Isa93i, Jan81b, Jon90a, Kau90, Kor19, Lan89g, Law85, Lee91b, Mec21f, Mec23c, MS77a, NA88f, NA88n, Pol79b, Pow90b, Ram86b, Ram10, RL78a, Rep89, Rot92e, Sim87b, Smi87a, Smi21, Tan84, TS22, Tsi77c, Ung85, Van20, Wol17, Wri85, Zys17, DeW81, Cra84b, Rot93j, Bie68, Gov93, Uhl98b, Bla87a]. Military-scientists [Bar81f, DeW81]. military-style [Ano91-40]. Milk [Goo00, Zag94b]. Millard [Sto59b]. Millennium [Hor75, Ano89-28, Ark98c, Moo97d, TD89]. Miller [Ano78d, Bri86, Lyn69, Pla57b, Abr04, Mar78a, Mec21a]. Million [Dow83f, Gre54, Ano13f, Ark96e, Rot93s, Spa95a, Wur93d]. Millions [Che52, Rab54a]. millionth [Ano90-33]. Millis [Joh51]. Millstone [Gor95b]. Milosz [Fer54b]. MILSTAR [Gor95b, Joh94a]. Milton [Ano88c, Cal84a, DeB81b, Kol63]. Mim [Ano89f]. Mimicking [Gan84]. Minatom [Jos02, Khr99a, Web02]. Mind [Bur05, Dum75, Fre67, Hac81, Hep77, Jor67, Lyo63, Opp49b, Sla76, Spe66, Fig23, Nei82b, Rot98b, Fer54b, Ker54]. Mind-Boggling [Hep77, Sla76]. mind-set [Fig23]. minded [LaB81]. minding [Wei97]. Minds [Eva04, For06, Rab62j, Eps12, Han95c, NA92o, Ray89, Smo89, Ede89]. mindscape [Ano12f]. mindset [ST22, Suz11]. Mine [Lew71a, Sku72, HC71, Moo97l, Wur93b, Wur93c, Wur94c]. Mined [Szi61b]. Minefield [Fly99b, Got82]. Mineral [Alb73, Ano46m, McK74, Sin76]. Minerals [Bat61, Mei51c, Mei51a]. Miners [Ano91-46, Mar89c, Cha82a]. MINERVA [Ell72, LEH^{+}71]. Mines [Rab98, Ven99a, Ven99b, Wur93a, Wur94b, Wur96c, Wur96b, Rei68b]. Mini [Opp04, NA92o, Ark94b]. Mini-nuke [Ark94b]. Mini-Nukes [Opp04]. Minibus [Ark94b]. minimal [Ble88, Szi64]. minimization [HNR12]. Minimum [Ing85, Lew08, Väy87, Wu60, Car83b, vH90a]. Mining [Ano55s, Dow82e, Fis46, MEA^{+}47, Sie06c, Sin81, Whi76a]. Minister [RI83, Tru83]. Ministry [Kir90, Voh71b]. Minneapolis [Col94]. Minnesota [Lap59a]. Minority [Eva54, FR76, Kol66, Sch05c, Isa87e]. Minute [Pai04f, Ark94d, Isa89h]. Minuteman [SA82]. Minutes [Ano74g, Ano02j, Boa07, Day74f, Dow82c, Jac74b, Rab50e, Bul04b, Sti83, Wei55b, Ack88h, Ack90a, Ano92q, Boa98, Fel80b, Fel81d, Fel84e, Hur88, Mas84, Moo98g, SM81, WM84, Cha79a, GR79, KFM50]. Minutiae [Wag75]. MIRACL [Low98]. Miracle [Rab47o]. Miracles [Rab52e, Zag95c]. mirage [Ano75g]. Mirages [NA87j]. Mirchandani [Gan83b]. Miriam [Tic94]. mirrors [Ano96j]. MIRV [Dow80e, De 70b, Gar67, Lap72]. misaligned [Kir17]. misbehavior [Tur18]. Mische [Ano78n]. Mischief [Sch60a]. Miscommunications [NA89o]. Misconceived [Ber95]. Misconceptions [MW77, Rob75a, Dav77a]. Misconduct [Fel83d]. Misdirected [Mat75, Nus80]. misfires [Ros89b]. Misguided [Wal81a]. mishandling [SM14]. mishaps [Aft86, NA89b]. misinformation [BV21, Sch24, YES^{+}21]. Misinformed [Jau90a, Jau90b]. Misinterpreted [NE76]. Misleading [Cam77, KE17, Sin91]. mismanagement [Smi82a]. misperceived [Len88]. Misplaced [Ark98e, Sho77, SD86]. misprinted [Ano92-34]. misread [Bur90a]. Misreading [Kul95, Uhl00]. miss [Ack90c, FD23, Shr23]. Missed [Ano92z, Sch00b, Sha95b, AH91m, Man86, Mil91, Tis90]. Misses [Stu48, Ada86a, Bea86a, Dro23h, Dro23e, Rev92, Rom90, Tin23]. Missile [Ack87d, Ano57l, Ano10a, Ben02, BBZ01, BS97, Bre67, Cle01, Cof65a, Cra67b, Dys69a, Dys72, Eva04, Gar67, Gar08, Han98a, Ing67, Ing68a, Isa98c, Isa99d, Isa01b, Isa04f, Joh01, Leb92, Leb94, Led71a, LP08, Loc91b, Mac86c, Mar67b, McM65, Men99, Moo02c, Mue01, Nol90, NA87h, NA89b, NA89s, NK12a, Pit91, Pod01, Pos00, Pug05, Rab67c, RC68, RD77, Sau01, Sel00, She00, She60, Stu89, Tsi75b, Ura01, WF77, Ani15, Ano87-38, Ano90-37, Ano18, Arb18, Bin18, BV10, Bru88, Cot19, Dre77, Ege16, Fie90, GN10, Gho11, GP18, Gre18, Gub15, Ike16, Isa89l, Kap83a, Kar17, KM89, KA91, Kre91b, KK19a, LvH18, Mec18b, Mey79, Nag15, Par87, Par17, RA67, Raj14, Riq15, Rob18, Rot93u, Rub90, SL13, Sch19, Sid16, Spa95d, Thr15, WW86]. missile [Zha15c, Kal68]. Missile-aneous [Loc91b]. Missileer [HW08]. Missiles [Bet69b, Dys64, Dys69d, Ful69, Haf47, HB63, Her90a, Iva02, Jac56, NA88l, NA90c, NK07b, NK09c, Pri58b, Rab57b, Rab58a, Rat69, Rol57, RC88, Sim57b, Ste69f, War59, Wol93, Yor69, You76c, Ack88l, Ano88p, Ano90-51, Ano96m, Bet89b, Cal84b, Cat87, FP90b, Gor95c, Har90b, Kar88c, Kar88d, KP98, Lor92, LA98, LGW92, Mar89f, Mil89a, NA87e, NA88h, NA89j, NA90m, Pow91a, Rev92, RK88, Sha84a, SLM87, Ste83b, The84, Tsi88, Tsi89b, War92b, Wol17, Ano87r, Hob85]. Missing [Ano92-27, Aud65, Bam85, Gor59, Kib04, Lap51b]. Mission [Bla62, Fer69a, Sub70, Sok01, Ano85y, Mel97, Oel12, WW86, Hac81]. Missionless [War92b]. missions [FP90b]. Missliski [Ric79]. Missouri [Ano84d, Kis93a]. Mist [Coo71]. Mistake [BCW74, Dys75, Aba89]. Mistaken [Mar89b]. mistakes [Dou24]. Mistress [Coo74, Mar73]. Mistrust [Bae60]. Misunderstanding [Sea71, CMR79]. Misunderstood [Min88, Wat73b, Car13]. Misuse [Ano69a, Arc54]. Misuses [Ark06]. Mitchell [Moo95e]. Mitchie [Ano15f]. Mitigate [Cav53, MR21]. Mitigating [KSS17, BV21, Eze16, Gra16, Wan16]. mitigation [FWCL20, Mil12b]. Mitra [Fry58]. Mitre [Rot96h]. mix [Ark86i, JP13]. Mixed [Kre86b, Pan98b, Fie91]. Mixing [KC03]. Miyako [Ano15f]. Mlad [AK98b]. MLF [Mar64e, Sha65, Sto66a, You64a]. mobiles [Arb91a]. Mobilization [Ano50z, Ano51s, Pas51]. Mobilizing [Bak75, Wad63, Hes78]. Moby [Wil52]. Mochovce [Wes98]. Mode [Rom90, BR85]. Model [Dal86, Whi71, BV21, Bru15, Dou15, Gre78, Sap88, Sha22, Vic78, War14, Ols74b]. modeling [Dro22i]. models [Hil12, Rob89, Ano81f]. Moderate [Mus00, Isa92a]. moderation [Nel94]. Modern [Ada52, Ada67, Ano87n, Aro64, Art68, Dav57b, Eps52, Fer57a, Ims60, Lar63, Mil85b, Nel93a, Pop52, Smi58f, Smi61, Spi72a, vL63, Dro22f, Hol76b, McM05, PMRP50, Ras55, Str67, Wri67]. Modernization [Ach12, Lin84a, Kro15, Lun82, McM82, Mec19e, Mec19f, Mia15b, Mou17, Tre19, WP19, Yin15, Loc91a]. modernizations [Kor17, KN14d]. Modes [Bak75, Dav69]. Modest [Gro06b, Dje76, Kre88e, Lin12]. Modification [Hol56, Jac73g, TT69]. Modify [CA98b]. modular [MR21]. moes [LSP95b]. Mohamed [Bul09f]. Mohawk [Sea56, Zim55]. Mojtabai [Gre87]. MOL [Smy64]. Mold [OG10]. Moldova [Gam94]. moldy [Ano98o]. Molecular [Nor06, Sac72a]. Molecules [Bea59]. Moloney [Ker54]. Mom [Rot94i]. Moment [Cir07a, Kin57, Mar02b, Cor92, Mar02a, Rap11, Sta97]. Momentous [Ano46-61]. Moments [Mar54f]. Momentum [JL74, Mil12a, Kib12, Nay12, Squ12]. Monday [Pla65]. Money [Aue04, Har88c, Isa02c, Lew61b, Rei02, Bul01b, Woo00f, Ano97m, Arn22, BV20, Dro21d, Moo95n, MRH16, RS15, Smi21, Spa95f, Ung85, Was97]. money-laundering [Arn22]. Mongolia [Poc05]. Monica [Ano98n]. monitor [YWS20]. Monitoring [AL62, Arb91a, Kup51, Len21, San63, Ano86w, JH16, KKS22, Dav79b]. Monitors [Pol62b, NA87j]. Monju [Tak17]. Monkeys [Her90a]. Monmouth [Ano54n, Ano55m, Mar54a]. Monod [Mar53c]. Monologue [Tur77]. Monopoly [Ler53, Net71, Goo84]. Monoxide [Ano73g]. Monsieur [Sta63]. monsoon [LG15]. monster [Wat77]. Monstrous [Cum94]. Montana [Jac73c, Ano79j]. Monte [Pen52]. Month [Rab56c, Isa85d]. Months [Spi57a]. Monticello [Gad72]. Monument [Col74]. Moo [Uhl98b]. mood [Kul95]. Moon [Arn73, Bro60a, Cow65, Lei73, Lew68c, RL69, Ure69a, O’N89b, Ric00b, And72, Ano63-27, Ano69m, Ano73e, Ano73f, Ano74h, Hym69a, Lew64c, Lov66, Mue69, O’K73, SJ68, Smi69b, SS73, Ure69b, Ure73, Vol74, dV61, Fol75, Lam69]. Moon-Doggle [Cow65]. Mooney [Har79b]. Moonquakes [Mid73]. Moore [Cha83b]. Moral [Arm53, Bel46, Chi78, Ein52a, Gol48, Hil46, Hol91, Hol66, Len76, Phi93, Wie56, Ein79b, Moo96k, Sha93c, Zac76, Mil85b, Fer53]. Morality [Bix49, Bur72, Doy17, Zei60, All86, Mac77a]. Morally [dL68]. Morals [SG64, Saw79]. Moratorium [Ano48-31, Dal75, Gla76e, Hur74, Isa94b, Jac74b, Cas77c, Gri83b, Wie82a]. moratoriums [Div86]. Moravcsik [Tel76]. Mordechai [Dow80i, BED93]. Morgan [Dow81g, Kat57a, Rob56]. Morgenstern [She60]. Morgenthau [How65b]. Mori [Bar90c]. Morison [Ada67, Saw80]. Morland [Gre82b]. Morley [Doh57]. Morning [Woo96c, Fel80g, Smi59a]. Morone [Per87b]. Morray [LaF62]. Morris [Bri86]. Morrow [Mor88]. Morse [Kat52a, Sal53, Wil77b]. mort [Bon57, Bon58a, Bon58b]. Mortality [Ano55u, Ano70b, NW70, Sag69, Ste69e, SK84, Tam69, SS24]. Mortgaging [Gre83a]. Morton [Ack87a, BGR61, Cof66, Ekl80, Mes85a, Rab64c]. Moruroa [Zam92d]. Moruroa-torium [Zam92d]. Mosbacher [Tho80g]. Moscow [Ano57l, Ano61i, Ano80a, Arb82, Ard93c, Bar83a, Bar81b, CR81, Cre90, Fra87, Hol91, Iff67, Iln97, Jac62a, Khr99b, Klo86, Les93e, Lon86b, Lon52c, Mil89b, Pai02, Pif18, Rab72e, Tro79, Ure45]. Mose [Fel73a]. Moskvie [Ano92x]. Mosley [Rep85a]. Mosquitoes [You62]. Moss [Fra82, Mar69]. Mossad [Mos91]. Most [Ano55y, Coh95b, DuB63, Gil83, Hud78, Lew62b, Rus53, Tod79, Dro24c, Har10, Per87a, RK19, Ric12, RW76]. mostly [Moo96n]. Mothball [Ano90-38]. Mother [And78b, Blo05, Isa89m]. motivate [HGF17]. Motivational [Str07d]. motive [TA94]. Motives [Dun49]. Motojima [Ano13f]. Motor [Ded62b]. Mott [Ano87-36, Ano87-37, Ano88r, Ano89x]. Mould [All59l]. Mountain [Ano46-90, Ano13c, CP98, CBSF95]. Mountains [Sag74]. Mounting [Ano74i, CSG71, Hau68, Rab71c, Smi72b, Bar77c]. Mouse [Pai81a, Ark96e, Arn22, Shu81, Swe81, Lan90d]. mouth [MH89b]. Mouths [Wad74]. Move [Bal76, Cha88b, Doc07, Lau60, Phi84b, STC74, Whi75, Yor69, Ano91-40, Ark84a, Bar10b, Bor87, Ekl77, ES89, Fel78i, Fel80b, Fel81d, HJS^{+}92, Kib10, Kor10, Lan10, Lyt85, Mat97a, Mec21b, Moo96a, NA90p, Pic10, Ram10, RL14, SS89]. Moved [Win98]. Movement [Bro47f, Hig66, Lor00b, Mor85a, Olu71, Pol91b, Rab69f, Sal69, Sch69d, Shi57, Swe86, Swe88, SM89b, Win94, Azh95b, BFG^{+}85, Cor92, Dav79a, Eve89, Mag16, Smi65, Smi70b, Wei87, Wil90, Win98, Vin59, Sch90]. Movements [Swe86, Wal83b, LS92]. Moves [Ano48n, Foz61f, Lov69, Asp87, Dea87, Isa89j, Lev89, Ust93a]. Movie [Rab51j, Dro23l]. movies [Ark86b]. Moving [Bou11, BHH11, Cam60, Hub60, Ing77, Mac96, Moo98f, RLG11, Rot93n, WC17, Ing76]. MOX [Cra98, Lev98, MvH97, Rep03]. Moxing [Ano97l]. Moynihan [Sch68e, Men91d]. Mozambique [Wur94a, Wur97]. Mr [Ada95, Cas77f]. Mr. [Ano47v, Ano49-31, Ano49-34, Ano52q, Ano53-41, Bun98, FDC52, GNMS54, Her60, Kih47f, Lev73, Mar48e, Rab47j, Rot97g, Rus58a, Saw53, Sha91b, Wyl57, A.52]. Msrg [Ano46-107]. Much [Ack87e, AM74, Dow83a, Ger75, Hun00, Lan90i, Mei51b, NAB00, Alb94d, Ano93s, Bar80c, Due21, Gar97c, Klu22, Moo93j, Moo95p, NA00d, Nuc18, PC92b, Rot98d, Smi79]. muddle [AK81]. Muddled [Deu91, GF79]. muddling [Fel85a]. Mueller [Boy89b]. Muj [Ste01]. mules [Gif13]. Mullen [Nor04]. Muller [Bal67, Car68]. Mulley [Wri63a]. Mulligan [Wol55]. Multilateral [New64, Par17, Sco89, SD15]. Multilateralism [SS05]. multinational [AS16, Hei11, Tat78]. Multinationals [Mei77b]. Multiple [Fin90]. Mumbling [Ber70b]. Mumford [Kat54b]. Munich [Fas66, Hib95b]. Munitions [Ark99e]. muove [Rab61d]. murder [Lag81, Moo93f]. Murphy [Bou98b, Kri85]. Murray [She61]. Murry [McC93]. Muscle [Hun00, Kla97b, Mer90, Rev92]. Musclebound [Wal99]. Museum [NA91i, Ano95x, Jr22a, HWG95]. mushroom [Ano83m, Tsi83i]. music [Rot99c]. Musical [FLLW91b, Rot92k]. Muskie [Mus65]. Muslim [Nas04b]. Must [Bar80m, Bli70, Boy10, Fed59a, Isa04d, Mos67, Myr74, Myr75, Pas77, Rab51k, Rab54a, Rab54b, Rab62f, Ram64, Rot04f, Sch54d, Sti68, Tam71a, Ure48d, Ano90-32, Con90a, Deu84b, Dro21a, Ekl77, Gei16a, Gou24, Ing76, Mec19c, Moo25, Pro85, Tru85, Wei76a, Wie57, Shu49, You66a]. Muste [Len88]. mutations [Hud76]. muted [KS98]. Mutual [Ede58, Lal77, Pla84, Sal78b, Wad58, Fel82e, Osn77, Väy87]. MX [Dow81g, You84a, Abt84, Asp83, Ben79a, Buc79, Bun83, GM87, Isa83c, Isa85b, Kin81a, Mey79, Orm85, Pai80b, Pai81b, Pai83g, Sno80, SP82, Ste85a, Wal81a]. MXed [Isa89i]. MXed-up [Isa89i]. My [Ano93x, Ano98n, Bor89, Fel83d, Pan07a, Pau52b, Wal69, AM21, Ano91-36, Ber23, Isa94f, JC54, Mae01, Sti18, Str56, Bor57b, Wil71, Bur65, Wil71]. Myanmar [Ahn10, ABKS10]. Myers [Gle94]. Myopic [Ack90i]. Myrdal [Ano86a, Tra71, Aur71d, Fel82g, Kam71, Myr72]. Myron [Pal64b, You66c]. Mysteries [Iln97]. mysterious [Dro23k]. Mysteriously [AHK90]. Mystery [Ano51t, Bru48, Kra78, Sie04d, FR98, OP98]. Myth [Ack87a, Act09, Bec71, Ber86b, Ded58, Hoo93, Iri56, Jon87, And81, Bro86e, Goe80, Kap78b, Kap78c, Mar97b, Mes91, MB86, NGG11, Wea85]. Mythic [FM93, Wil68a]. Mythical [Gou70, Gra93]. Mythology [Cha64, JL74, Lil63, Buc81, Chi82, BCB81]. Myths [Bul83, E.77, Gan83b, Gol69a, HSJH18, Rom90, Tel82, Van80, Wil78b, vH83a, vH83b, FS81, HCH16, IK21, You87a].

N [Bou61, Ekl80, Fin90, Fla90, For80, Kat52b, McC96, Mil86b, Pat81a, Sch63c, Sto59b, Ano91-36]. n-word [Ano91-36]. N. [Ano48s]. N.S. [Ano59e]. N.Y. [Ano46-68]. Nagai [Amr51]. Nagasaki [Cog82, Bal54, Bar77b, Ber75b, Dow80h, Dow82h, Dow83b, Fel66b, Fel85b, Iwa82, Loe85, Loe95, Nat77, NA89s, Osg85, Ric12, Amr51, Bru57, Bru59, She83a]. Nahum [vH77a]. Nairn [Ada61]. Nairobi [Spu80]. Nakaya [Smi54b]. Nakayama [Sak69]. Nakdimon [Ram88b]. Namboodri [Gan83b]. Name [Ano91-40, Dal75, Spi66a, Ano82l, Ano90-30, Ano98o, Bon90, DM82, Hor88c, Isa93g, Pol91b, She80]. Named [Ano46-62, Ano49u, Ano75a, Ano85l, Lew68a, Lut73, Ano97i]. Names [Ano62t, Ano62-28, Kih47f, MT03]. Namibia [Mac81a, Mac81a]. Nanda [Fer53]. Nanking [Kre98]. Nanomaterial [McN09]. nanotechnology [Kos10, Kos19, NF13, Tre07]. nario [Alb94b]. Narrating [Gus09]. narrative [Mor83c, Squ21b, Tsi83c]. Narrow [Dou68a, Rab48i, Vie62, Vol74]. Narrowing [McC54b, Rab53e]. NAS [Moo97o, Rot92n]. NASA [Ano93z, Fru61, Lew70b, Man86, New61, O’L73, Bro67a]. Nash [Dow83b]. Nasr [Ano15g]. Natasha [Ano94k]. Nathan [Kap61, Nat68a]. Nation [Ada87, Ano50i, BGR61, Cha76, For51a, For51b, Lea07, Shu78, Tho80g, Ano90-27, Coh95b, Fer54a, Hel85, Hol91, Kre81, RK19, Bar48, Don86b, Lew62a]. National [Adv60b, Ano46-73, Ano48h, Ano51p, Ano78k, Ano81f, Ano84d, Ano88e, Ano89d, Aug48c, Bar60, BW59, Ber53b, Bra47, Col74, Dem60, Dad71, Dan48, Dea52, Deu83b, DuB46b, Duf70, Dum76, Fel77o, Gar72, Gla56a, Gre72, Gro51, GR63, Gus11, H.84a, H.84b, Ham93, Her83a, Hil60, Hir55, Ing56a, Isa00c, Isa01c, Jav68, JL74, Kri14, Lap61a, McC85, McN64, Mec21c, Mes64b, Moo92h, Mor70a, Mor70b, Mor62d, Mye87a, Nel84, NA01, Pai82h, Par53, Par46b, Par47, Pit64, Pos00, Pot53a, Pot53b, Rod69, RR81b, Rub08, Stu47, Sar78, See79, Sel00, She83b, She60, Spi66b, Spi67, Sta00, Ste89a, Ste46, Tay98, Tir88, Tru51, Tuv53, Was64, Wat53, Wat57, Whe51, Win71, Woo82, Yos80, You66a, Zac50, vL92, BBL13, BB79, Cha82c, Ded78, Den14, Fla80a, Hal85]. national [Isa93e, Ken63c, Kir17, Lan90l, Ode88, Sch21, Sha78a, Squ13, Sta81, Tur13, Wei78d, Ano45h, Ano45m, Ano46-63, Ano46-62, Ano46-94, Ano47o, Ano47p, Ano50-27, Ano52l, Ano52s, Ano53x, Cha88b, Dub50a, Fro49, Gor51b, Gro52, Gro53c, Han69, Har79a, Lac85, May47a, Mey49b, Osb58, Pri47, Stu48, Sci49, Cou54, Smy50b, Tru50, Wat51, WG69, Ack97, Sha81a, She82, Sla88, Yon68]. Nationalism [Bat46b, Ich64, Sch71b, Wil68a, NA88h]. Nationalities [Kol66]. Nationality [Ano55b]. Nationalization [McC69]. Nations [Ano48-39, Ano55r, Ano55y, Ano78n, Ano78m, Ano90o, Apt61, Bar82a, Bél48, Dav07, Dje68, Eis54c, Har87d, Hes78, Hud78, Jan64, Jon59, Lal70, Lon69b, Myr75, Nas62, Noy46b, RE68, Sal64b, Sal73, Sto58, Upt63, Bak78, FP17, Lop78, Pai85d, Pol76, TT83, Van91, Ano46-105, Ano47c, Ano47y, Ano51-28, Ano52-27, AE85, Bec75, Car49b, Car49c, Car49d, Car49e, Car50a, Car50b, Car50c, Car50d, Car50e, Car50f, Cav56, Cor78, Eps82a, Eps82b, Eps92b, Eps97, Fry57a, Hud76, Kih47c, Kih47d, Kih47f, Kih47e, Kih48a, Kih48b, Kih48c, Kih48d, Kih48e, Kih48f, Kih48g, Kih48h, Kih48i, Kih48j, Kih49, Llo71, Osb49a, Pai82f, Pay70, Pea85, Pea91, Pea99, Sal68, Ste52, Wri61, Wur92, dC78, Ded62a, Men91d, Nad98a, Nas62, Rab67a, Wri47a, Lal65c]. NATO [Ler80a, Ano87-31, Ano91-34, Ano92t, Ark86a, Bor89, CAM20, CA84, Cha85b, Cha87, Coe98, Dea96, Eis59, Gol94, Gor59, Hat85, Hay81, Isa97d, Isa98a, Iva02, Kom90, Lal64e, Lüb90, MB98, MM98, Mei06, MH89a, Mot82, Nel92a, Nel94, NA90h, O’H01, Pai82g, Pav98a, Ple91, Pow98, RK89b, Sau17, SS89, Sei60, Sha85c, Sha87d, Sha92b, Sha95c, Sha98c, Sho92c, Sig89, Swe86, Woo99b, Cro65c, Nie65, Lin84a, Sei60, Wri63a, You66b]. Natural [Ano78n, Gin54, Smi54b, Wil73c, Fre77b, Pat90b, Wol78, Noy47]. Nature [Adr54, Ano46-40, Ano70e, Bea64, Bla84, Cro67, Ehr81, Fer57c, Har70, Mar64j, Mar58d, McC93, Pla65, Tel75, Tel57b, Tho80g, Woo65, You82, Ano97n, Ano98k, Ano15g, BMPH18, Dav77a, Tho80g, Ehr82a, Haw50, Lan57, Sch86, Tod81, Ano46y]. Naval [Ano46-71, Ano87q, Ano90p, EB73a, NA93i, WC47, Hun90, Moo15, NA88c]. Navies [Sid00]. navy [Han93, NA89f, Ano46k, Ano46-64, Ano56h, Ano87-38, Ano91-41, Ark87a, Ark87b, Bro47a, Bro47b, Dow81f, Flo89a, GM87, Mae01, McM54, Mcc87, Moo15, NA87g, NA89b, Ros75, SM14, WM78, Zim89b, Pap88, Pla60]. Nawapa [Ano67i, Cru67, Lew65b, Ost67, Roy67, Sew67, SSC68, Tin67]. nay [Ano89i]. nays [Isa88c]. Naysayer [Mue07a]. Nazi [Kor92, DR92, GP92b, Hun85b, OH85]. Neal [Dow83c, Smi60, Pat79a]. Near [Dro23h, Leb94, Moo03b, Van61, Ano59d, Cum98, Day77h, Dro23e, Far98, FD23, Gho11, Gor96d, Hum83, Nel12, Shr23, Soc15, Tin23, Zan91]. Near-death [Leb94]. near-Earth [Far98]. near-miss [FD23, Shr23].

Near-misses [Dro23h, Dro23e, Tin23]. near-term [Nel12]. near-war [Cum98]. Nears [Kre99, Sig88b]. Necessary [Fla60, Gra82b, HKC78, Pra03a, Sto03, Ano86s, BW80, BBG^{+}85, HS13, Pan86b, Ste77c, Tho80a, TKG94, Wei80a]. necessity [Ber86c, Har76b, Lin12]. neck [Joh94a]. Ned [Bla84, Ble87, Kra94, You82]. Need [Ano80k, Ant74, AM74, Bet77b, Civ02, De 74, Ehr72, Fed61, Kal70, Ken61c, La 66, Lor74, Mas57, May52, Noy57, Osb54, Ram64, Ram70, Rod69, Sak75, Sap75a, Sea68, Smi75d, Sta49, Bul08b, Ano79k, Ano93u, Ano96n, Gev93, GG11a, Gre12, Kel10, Lal77, MFN17, MRH16, Mül90b, NH08b, Rap11, RL76, Sea89b, Sin92, Woo92, vH16]. Needed [Buk03, Can74, Com63a, Day74c, Eis99, Len60, Rab72f, Rom91, Sak98, Sol76, Spo62, Ste58, Str78, Ure49a, Ano90z, Bun89, Fig23, Kau80, Moo14]. needles [Kum95]. Needs [Ano78g, Bal79b, Bla67, BM08, Fit51, Lew61b, Not51, Ste75b, vL48b, Abb12, Bun82, Cam90, DJ13, Gar78a, Ger17, Kan96, Kor21, KSG15, Man21, Mül90a, Par17, Rei91, Sal88, Sha96, Smi21, Tho88]. Needy [Kha99b, Kha99a]. Neel [Bru57]. Negating [Wil95]. Negative [Har77b]. Negatives [Voh70]. Neglect [Fre02, Gar92]. Neglected [Rex71, Tho75, GK85]. negotiate [Dro23g, Far87, Swe79b]. Negotiated [FGJ65]. Negotiating [Bun89, Cra84b, Kin86, Kwo10, Men92b, Mor50c, Pan07a, Sig03, Yor83, vMK88, Div79, Kin86]. Negotiation [Mar55b, Wri67, Cra84b]. Negotiations [Ano60r, Bet54, Bur09, Cal73, Cav61, Dav60, Fre83, Gol09, Hum59a, Kin86, Kre93, Lal64a, Mar48c, Tur60, You61, Gra82b, Gro86, Har85a, Har86b, Har86c, Har87a, Har87b, Har88a, Kis79a, Lin84b, Lal67b]. negotiator [Obu89, Riv96]. negotiators [Has85]. Neighbor [Pug05, Ano95n, Ram88a]. Neighborly [Ano98l]. Neighbors [Guz52, Moo92i, Mos52, Val52, Woo98f, dO52, Ano93r]. Neil [Gri83a]. Neill [Bar67a]. Neinaber [Yar80a]. Neither [Her95b, Kri92b, Mei01, Ano89i, Kae53]. Nelkin [Sim84b, Wal83b, dF80, Nor88, dF80]. Nemesis [Hag77]. Nemet [Dro21a]. Neo [Blo71, Tsi78b]. Neo-Isolationism [Blo71]. neo-Luddite [Tsi78b]. Neocons [Hus03]. Neoconservative [Kre86c]. Neophytes [AB08]. Nepali [Gya14]. Neprípustné [Jan76b]. Neptune [Han90d]. Nerds [TB23]. Nerlich [Cra84b]. Nerve [Pay84, Ano60i, Isa83a]. Nervous [Isa83a, LLH^{+}62]. net [Mat98f, Utr87]. Neta [Ano88d]. Nether [Ikl75]. Netherlands [Ano80g, Min48, Myn11, SOE21, SS76]. Netting [Ano94j]. Netwar [Zit17]. Network [Sai09, Sap68, She05b, Bye91, Fos88, Fre91d]. networks [Ban11]. Neuhaus [Bre73]. Neumann [Tel57a]. Neuneck [Ano10a]. Neuro [Fav52]. Neuro-Surgery [Fav52]. Neurosis [Hym07]. Neutral [Eme61, TT83]. Neutrality [Gil50, Lie62a, MW75, Tiu94]. Neutralization [Lal67a]. Neutralize [Bro61]. Neutrals [Lal67b]. Neutrino [Dan52]. Neutron [All59l, Ano61o, Dys61, Eme61, Lan98, Mor83b, Olk77, Rob78c, TFM81, Ano91-33, Dys20, Fel77m, Fel81j, GBG^{+}81, Kap81b, Kis78, Oli82, Sør77]. neutrons [TB23]. Nevada [Bor53, Gea89, MD64, NA89f, Tay91, Zei52, Zhe90b]. Nevadans [Rot92m]. Never [Ano96m, Ark01d, Bar80m, Coh00, Fre02, Moo95m, Rei02, Alb94e, Ano95l, Cli77, Cum97b, Day76e, Hec22, Kno90, Lei89a, RR88]. never-to-be-developed [Cli77]. Nevil [Doh58]. New-Nuke [Sch01b]. Newberg [Woo82]. Newbomb [Ric79]. Newcombe [Lyt79]. newcomer [Al 11]. Newly [Pre60, Hes78]. newly-formed [Hes78]. News [AH91m, All54c, All55c, All55d, All55e, All55f, All55g, All55h, All55i, All55j, All55k, All56a, All56b, All56c, All56d, All56e, All56f, All56g, All56h, All56i, All56j, All57a, All57b, All57c, All57d, All57e, All57f, All57g, All57h, All57i, All57j, All58b, All58c, All58d, All58e, All58f, All58g, All58h, All58i, All58j, All58k, All59b, All59c, All59d, All59e, All59f, All59g, All59h, All59i, All59j, All59k, All60a, All60b, All60c, All60d, All60e, All60f, Ano45e, Ano45m, Ano46-31, Ano46-32, Ano47c, Ano47b, Ano48l, Ano48m, Ano48n, Ano50j, Ano50k, Ano50-35, Ano50-30, Ano50-31, Ano50-32, Ano50-33, Ano50-34, Ano51e, Ano51f, Ano51g, Ano51t, Ano51-28, Ano52n, Ano52o, Ano52p, Ano52-27, Ano53z, Ano53-27, Ano53-28, Ano53-29, Ano53-30, Ano53-31, Ano53-32, Ano53-33, Ano53-34]. News [Ano53-35, Ano54l, Ano54m, Ano54r, Ano54s, Ano54t, Ano54u, Ano54v, Ano61o, Ano61p, Ano61t, Ano62s, Ano62t, Ano62-28, Ano63u, Ano63r, Ano63s, Ano63t, Axe50, Axe51k, Axe51h, Axe51i, Axe51j, Axe51g, Axe51b, Axe51c, Axe51d, Axe51e, Axe51f, Axe52a, Axe52b, Axe52c, Axe52d, Axe52e, Bea47b, Bro62a, Cal63, Car49b, Car49c, Car49d, Car49e, Car50a, Car50b, Car50c, Car50d, Car50e, Car50f, Dux52, Foz61c, Foz61e, Foz61d, Foz61g, Foz62b, Foz62c, Foz62d, Foz62e, Gol62b, Gre62, Ham63, Hin69a, Hut50, Hut51, Jaf51a, Jaf51b, Kih47a, Kih47b, Kih47c, Kih47d, Kih47f, Kih47e, Kih48a, Kih48b, Kih48c, Kih48d, Kih48e, Kih48f, Kih48g, Kih48h, Kih48i, Kih48j, Kih49, Lal64f, Lev52, Lew63b, Mar60a, Mar60b, Mar60c, MF63, New62b, Pre63, Rab47n, Rab51l, Sch63a, Sim54c, Sim54d, Sim54e, Sle52]. News [Sle53a, Sle53b, Sle53c, Sle53d, Sle53e, Sle53f, Sle53g, Sle54a, Sle54b, Sle54c, Sle54d, Som83b, Ste52, Tru46, Wid60, Wid61c, Wid61d, Wid61e, Wid61f, Wid61g, Ano88o, Ano90-47, Ark95e, Fel77d, Gol90d, Mat98f, Mec18a, NA90q, Rot96e, Spe85b, Sto18c, Tat78, Ano48f, Mec21a, Ano46-56, Ano54y]. newsletter [Dro23m]. Newspaper [Ano48e, Szi61a]. Next [Ano63t, Ano90-39, Ano98l, Bas80, Cum03, Dam62, Deu67, Dia83, Eps96a, Fel68a, Fen91, Fer10, Gre54, Isa00c, KK02, Lal66g, Pie54, Pla57b, Pot70, Rab51j, Rob69, SA62, Shi48c, Smi70c, Sto97, Sto66a, Sto67, Wal84, Ach10, AO95a, Ano89-27, Ano90-32, Ano12g, BL83b, Cro18, Din78, Dro20b, Jr22b, Dro24f, Fel80h, Hic22, Hor14, KSS17, Küh17, Lin84c, Mag16, McL11, NH08b, NA94e, Nye77, Oel12, Sch16, Set16, Sha92a, Sim16, Soc20, Sok16, Spe87a, Wat19, Zal19, Sco72]. next-generation [Hic22]. Nexus [CKA10]. NGOs [Cle99]. Nguyen [Mar94]. Nicaraguan [Law90]. Nice [O’H01, Sig85b]. Nicholas [Ano56c, Bra89, Eli85, Kat57b]. Nicholls [Kin69]. Nichols [Pee78, GNMS54, Nic54b, Nic23]. nickname [Rot95e]. Niebuhr [SD55].

Nieburg [Spi66a]. Niedringhaus [Wil80b]. Niels [Opp63d, Pet63, She86b, Wil85b, All56l]. NIF [Gus95]. NIF-ty [Gus95]. NIFty [Ano95j, Mat97a]. Nigel [Fel68b]. Nigeria [Kon67, Sin05a]. Nigerian [Fal61b, Eze16, Tom14]. Night [Cha06a]. Nightmare [Ano99e, Elk62, Ram05, Sie05f, AG98, Kam96]. Nightmares [Bar80m, ML10]. NIH [Nov77b]. Nike [NA89f]. Nikita [Leb91b, Rit67, WS21]. Nikitin [Mat98a, Han00a]. Nikolai [Ekl80]. Nikolaii [Ekl80]. NIMBY [Shk92]. Nimkoff [Pla57b]. Nina [Bal79c, Fin65]. Nine [Boa98, Bal54, CCD^{+}52, Moo98g, Ano93-27, Loe88, Bru59]. Nineties [Hob85]. Ninety [Isa83b]. Ninety-eighth [Isa83b]. ninjas [TB23]. Nintendo [Cha88b]. Ninth [Bal79b, Bos51, Tur51, Say84]. Nippon [Ano91-47]. Nitze [Kre90, Cah94]. Niu [Ano92x]. Nixing [HT03]. Nixon [Gal86, ADH^{+}73, Ant74, BK03, Kar72, Lon73, Mes70, Tay74]. NJ [Ano48d]. NL [BM99]. NM [Ano22c]. NMD [Gro01]. No [AC76, Alv74, Alv03, Alv20, Ano46m, Ano46-67, Ano79k, Ano93s, Ano93u, Ano93-27, Ano94m, Ano94l, Ano95m, Ano96n, Ano97i, Ano98l, Ano98o, Ark86e, Ark98f, Ark00b, AS94, Aue05a, Bar64a, Ber04a, Bro88b, Bru96, Car83c, Cav92, Cha76, Coh02, Col88, Coo58, Day75i, Dea96, Dou60, Dys75, Eba72, Ein50, FR97b, Fel67d, Fel76e, Fel79g, Fla60, Fly00b, Fre77a, Gir75, Gla63, Goo77, Gor96b, Gor96f, Isa89l, Isa04c, Jac71a, Jac74b, Joh01, Joh04, Joh05, Jon97, Kib03, KR84, Kle81, KR12, Kri85, Len73, Len76, Lew67a, LN83, Loe89, Lon73, Mar89c, Mar93b, Mar83e, Mas57, Mat97d, Mes83a, MM90, Mil20, Moo94h, Mot48, Mot83, Mül90b, NA90l, PWS^{+}96, PS04, Plo86, Pra96, Pra04b, Pri89, Pug05, R.83, Rav83, RMW54, Rei67, Rob91b, Rom77, Rot92n, Rot92o, Rot92b]. No [Rot96c, Rus92a, Sap84, Sas06, Smi74b, Smi85c, Smi96, Smi92b, Sør77, Bul01b, Sto03, SSR83, Str06c, Str94, Swe02, Uhl95a, Uhl01a, Ung78b, Web06, Wei83b, Wie78, Wil75a, You67a, Zag95c, Zar76, Zim89b, Ano90-31, Ano91-34, Ano95l, AH89a, Ark97d, Ark97g, Asp78, Ber01, Cle92, EP86, Fel80e, Hol20, Hug95, Hyd91, Isa89m, Isa92n, Isa94h, Jon91a, Jon93a, Kan96, Kap92, KB80, Kid91, Kis93b, Lan89f, Lut89, Mac92, Mel97, Moo96n, Nel92a, NA89s, Ola82, Pie19, Rom91, Rot77, Sco95, Sol90, Toy95, Vog90a, Wan96, Weg96b, Wil15, Rab73g, Arc86, Bet75a, BAA^{+}75, Dow80f, Gar88a, Gar88c, Gro83, Han93, Lan79, Lin86, Mar90, Nit80, TC83, ZG83, dF79h, vHWF^{+}81, BK81, Bun82, Ano83k, Beu49, Dow83e, Dow83d, Gol69b, Ham83, Kri85, Mor61a, Sha59]. No-brainer [Moo94h]. No-Dual-Use [Ber04a]. No-First-Use [Len73]. No-LOW [PS04]. no-nuke [Cle92]. no-regrets [Rom91]. Noam [Cro17d]. Nobel [Kat52b, Nad79, Tis52, Ang78, Ano85m, Ano87-32, Bal79a, Bun79, CCD^{+}52, Day77i, Day96, Dro16a, Dro23a, Fel82g, Fra76a, Kas75, Mec20f, Nel77, NB79, Orr78, Pau79, Sta79, Sut97, vS78, Kat52b]. Noble [Ano90-40]. Nobody [Ano94m, HD82, Rus58b, Rus20]. Nobre [Dro25a]. Nod [Fer99]. NOEMs [Bal87]. Nogee [Ham61b]. Noise [Ant69]. Nolan [Mec23a, Tra90]. Nomenklatura [Lon85c]. NOMOR [Sol77]. Nomos [Ano10a]. Non [Alb93e, Cal60b, CG04b, DeV62, ES89, Fra83, Fre73, FC96, Jac63b, Kin74, Lof60, McD13, Moo01c, Que69, Rab69g, Rod61, Sal75a, San90, Tod64a, Wei78e, Aft94a, Aue03d, Eps96b, Fel80l, Fel81b, Gra81b, KM18, Lan82b, Les81, Mog81, Odh81, Sap80a, Sau17, Tan84, Wei18, AC85, Ano94s, Don86b, Eps81, Hal75b, Ima69, Lev85, Ros94]. Non-Aggressive [Moo01c]. Non-Aligned [Jac63b, Sal75a, Eps96b]. Non-Apocalyptic [Tod64a]. non-economist [Sap80a]. Non-Governmental [Rod61]. non-issues [Fel81b, Gra81b]. non-lethal [Aft94a, Aue03d, Ano94s, Ros94]. Non-Members [CG04b]. non-military [Tan84]. Non-Nuclear [Cal60b, DeV62, ES89, Sau17]. Non-P-5 [McD13]. non-policy [Lan82b]. Non-Political [Rab69g]. Non-Polluting [Fre73]. Non-Proliferation [Fra83, Que69, Wei78e, Alb93e, San90, Fel80l, Les81, Mog81, Wei18, AC85, Don86b, Eps81, Hal75b, Ima69, Lev85]. Non-Retaliation [Kin74]. Non-Scientific [Lof60]. non-traditional [KM18]. non-use [Odh81]. nonaggressive [For88]. Nonaligned [Jac64b, Lal64a]. nonalignment [Lam85]. None [Dow81g, MW47, Win94, Cha75d, NA90p, Ano46u, WM82]. Nongovernmental [Bou85]. Nonimmigration [Ano55o]. Noninterference [Sti68]. Nonintervention [Ano66s, Lal66d]. Nonmilitary [Bie68, IS67, Par53]. Nonnuclear [Cro64, Pod16]. Nonoffensive [Ano88o, Fla88, KWM88, Kap89]. Nonpartisan [Mes84a]. Nonprofit [Wei53a]. Nonproliferation [AO95b, AO95c, Ano66n, Bai95, Coh95b, DR71, Fel64, Fel68c, Fra83, Gol91c, HB10, Kap95, Kra92, Que67, RDS85, Sai09, Sie04c, Sim90a, Sim90b, Bel20b, Ege16, Ike16, KH13, Lan89a, LP85b, Low85, Mec20b, Nat85a, PB92, Sha10, Sid16, SS90a, Fel67a, Fel68a, Har07, Joh68, Ham93]. Nonsense [Bos78, Bel94, Bro85b, Car82b, Cat89, Isa88e]. Nonstrategic [KN12b, Zys17]. nonviolence [Ano83i]. noodle [Say92]. Noon [Hin60]. Nor [Mei01, Ano89i, Her95b, Kri92b]. Norden [Kap61, Sto24]. Nordic [Gri85, ML82]. Norie [vH77a]. Norm [Wad74, Sha10]. Normal [Mel75]. Normalcy [Rab54g]. normality [Mas13]. Normalizing [Pea84, Bre84]. Norman [Fra82, Gle94, Lyt79, Mar69, Mar54f, Sto59b, Wou81, Wri47a, Rou75]. norms [Kor16]. Norris [Ano14a, Han94]. north [Ark84a, Dje76, NA91g, Ahn10, AH91n, AH92h, Alb93e, Alb94b, Alb94d, AH02, ABKS10, Alv03, Alv20, Bin10, Cum15, Eav17a, G.65, Gil98, Gol09, GP18, HSV19, Hay93, Hea83, Hec08, HC12, Hol85a, Hym07, Kan09, KN18b, KK21b, KK22c, KKJK24b, Kwo10, Lan16, LST10, Man80, Mat82a, iM16, NKH03b, NK05d, Pan22, PPE22, Sam18, San91, She16, Sig01, Sig03, Sig10, Bul03b, Wat19, Wri99a]. north-south [Dje76]. Northeast [Jac80b]. Northern [Ano46e, Cum06, Opp02a]. Northrop [Isa91j]. Norton [Nei82a]. Norway [Ano62-28, Ano63w, Gle83, Ran50]. Norwegian [Fel78e, Ran53]. nose [Abb93]. nosed [Fel77e]. nostalgia [Boy05]. Not-So-Selective [Ahm03b]. Note [Ano52q, Ano53e, Ano53o, Ano54-29, Ano66g, Ano66h, Ano66i, Ano67e, Ano73g, Ano74j, Ano75e, Ano85b, Ano99f, Ano01a, Ano16b, Coc51, Cra83a, Fel66a, Fel84b, G.84a, Gan83a, Gan84, Gri83a, H.84a, H.85, vKJ69, Lew72b, M.85, McG84, Mec16a, Moo97f, Moo97e, Nel84, Rab59b, Rab60e, Ach56, Ack88b, Ano79e, Ano86h, Ano87m, Ano87l, Ano88d, Ano88c, Ano89e, Ano89d, Ano90n, Ano90m, Ano91q, Ano91r, Ano93n, Ano94h, Ano98g, Cra82b, Cra82c, Cra83d, Cra84b, Cro16a, Cro16b, Cro16c, Cro16d, Cro17c, Cro17a, Cro17b, Day77e, Fel77k, Fel78k, Fel78l, Fel78m, Fel80k, Gen84a, Gra84b, Mec20d, Moo92b, Moo92e, Moo92d, Moo92c, Moo92f, Moo92g, Moo92i, Moo92h, Moo92j, Moo93c, Moo93d, Moo93e, Moo93f, Moo93g, Moo93h, Moo93i, Moo93j, Moo93k, Moo94c, Moo94d, Moo94e, Moo94f, Moo94g, Moo94h, Moo95f]. note [Moo95h, Moo95g, Moo95i, Moo95j, Moo95k, Moo96k, Moo96h, Moo96i, Moo96g, Moo96j, Moo96f, Moo97g, Moo97h, Moo97j, Moo97i, Moo98d, Moo98e, Moo98a, Moo98f, Moo98g, Moo98h, Mye93a, The88, The89a, The90, Tsi83h, Tsi83g, Tsi84b, Wei78b, You84b, Ano55p, Cra82d, Fel82h, Gro83, Rev83, Tsi83d, Tsi83e]. Notebook [NA87e, NA87f, NA87g, NA87i, NA97c, NK09b, NA87c, NA87d, NA87h, NA87j, NA88b, NA88c, NA88d, NA88e, NA88f, NA88g, NA88i, NA88h, NA88j, NA88l, NA88m, NA88n, NA88o, NA88k, NA88s, NA88a, NA88r, NA88t, NA88q, NA88p, NA89b, NA89c, NA89d, NA89e, NA89f, NA89h, NA89g, NA89i, NA89j, NA89a, NA89k, NA89l, NA89m, NA89n, NA89o, NA89q, NA89p, NA89r, NA89s, NA90b, NA90c, NA90d, NA90e, NA90f, NA90g, NA90i, NA90h, NA90j, NA90k, NA90l, NA90m, NA90o, NA90n, NA90p, NA90q, NA90s, NA90r, NA90t, NA91b, NA91c, NA91d, NA91e, NA91f, NA91g, NA91h, NA91i, NA91l, NA91j, NA91k, NA91m, NA91n, NA91p, NA91o, NA91q, NA92a, NA92b, NA92c, NA92e, NA92d, NA92f, NA92g, NA92k, NA92i, NA92h, NA92j, NA92l, NA92m, NA93a]. notebook [NA93b, NA93d, NA93c, NA93e, NA93g, NA93h, NA93k, NA93j, NA93i, NA93l, NA93n, NA93o, NA93m, NA93f, NA93p, NA93r, NA93s, NA93q, Nor94a, NA94a, NA94b, NA94f, NA94e, NA94c, NA94d, NA94h, NA94g, NA95b, NA95c, NA95a, NA95d, NA95e, NA96f, NA96d, NA96c, NA96e, NA96b, NA96a, NA97e, NA97a, NA97d, NA97f, NA97b, NA98g, NA98d, NA98e, NA98f, NA98b, NA98c]. Noted [McR51]. Notes [Ano48n, Ano52n, Ano52o, Ano52p, Ano53z, Ano53-27, Ano53-28, Ano53-29, Ano53-30, Ano53-31, Ano53-32, Ano53-33, Ano53-34, Ano53-35, Ano54r, Ano59g, Ano59h, Ano61d, Ano62z, Ano85d, Ano85c, Ano87n, Ano90o, Aur71e, Axe51g, Axe51b, Axe51c, Axe51d, Axe51e, Axe51f, Axe52a, Axe52b, Axe52c, Axe52d, Axe52e, Bou66, Clo71, Dux52, Ehr72, Gom75a, Gom75b, Gom75c, Gro53d, Hah67a, Hah67b, Jac71b, Lal65e, Les93c, Mar64c, Mar64d, Mar64e, Mar65f, Mar65g, Mar65c, Mar66a, Mar49f, Mar49c, Mar49d, Mar49e, Mar51, Mar53a, Mar53b, Mue07a, Mul73b, Rab51l, Sim54c, Sim54d, Bul05, Ano86k, Ano86l, Ano86j, Ano86i, Ano87p, Ano87r, Ano87o, Ano87q, Ano88e, Ano89f, Ano89g, Ano90p, Ano91s, Ber79, Sim81, For80]. Nothing [Ano75f, Ano95l, Com75a, Kis93b, Rot95k, Bra13b]. Notice [Ano92-34]. notices [Ano92t]. Notion [Lev07]. Nouveaux [Ano96p]. Nov [Cor53]. Nova [Can79, AL79a]. novel [Mag81, McC88a]. November [Ano52b, Ano61i, Ano63q, Chu54b, NK12a, Ano48u, Ano53q, Ano83h, Fel80g, Isa85d, Kal82e, Mos79]. Novick [Ing70, Lon78]. Nowhere [Coh02, Rab56m].

Noyes [Ing79a]. NPG [Mul74b]. NPT [Ach10, AZ91, Ano90-55, Ano90-60, Ano90-59, Bar80c, Bar80k, CB95, CG04b, Dha90, Dun90, Eps75b, Eps90b, Eps90c, Eps95, Fel80l, Fis90, Gol86a, Gol90a, Joh99b, Joh00a, Joh10, KN14d, Loe86, Mia98, MF68, Muk93, Mül85, Mül90a, Nat86b, NA93j, Rab69f, Rot04c, San90, San91, Sim90b, Sim91, Sok20, SS90a, Spi74, SK03, Sub93, Tho90, Wei17, Wil95]. NRA [Eis90, Gol99]. NRC [Ano12g, BS12, Hir02, HLL03, Kou75, Lan79, vH80]. NRDC [Ano86w, NA95a, NA96f, NA96d, NA96c, NA96e, NA96b, NA96a, NA97e, NA97c, NA97a, NA97d, NA97f, NA97b, NA98g, NA98d, NA98e, NA98f, NA98b, NA98c]. NSC [Tod79, vL92]. NSF [Ano51u, Gar68]. NSRB [Ano51v, Zei50a]. Nuckolls [Wei94e]. Nucléaire [Dav94]. Nuclear [A.55, Adv60b, Abb85, Abb89, Abb91, Abr73, Ack87a, Ack90j, Ack94, Ack09, AB08, Aft86, Ahm99, Aik61, bZAAFC08, AH91n, AH92l, AG98, All54d, Alm85, Alv05, And85, Ano46-68, Ano46-69, Ano46-73, Ano48p, Ano54u, Ano55g, Ano56i, Ano58k, Ano59i, Ano59j, Ano60d, Ano60f, Ano60k, Ano60o, Ano61k, Ano62p, Ano62v, Ano65g, Ano67d, Ano78f, Ano79e, Ano80q, Ano81f, Ano84g, Ano85d, Ano85c, Ano85n, Ano86k, Ano87p, Ano87q, Ano87s, Ano90-41, Ano91s, Ano91r, Ano91-66, Ano93l, Ano93-33, Ano96n, Ano99m, Ano99o, Ano99-27, Ano00o, Ano01b, Ano02i, Ano10a, AM86b, Ark83a, Ark83b, ACH84, AF85, Ark85d, Ark86f, AH89b, Ark89c, Ark91, ADK91, Ark93a, Ark96b, Ark96f, Ark00c, Ark00d, Ark01a, Arm68, Arn96a, Arn53, AB05, Aue02, AC86, Bac49, Bac77a, Bar77a]. Nuclear [Bar66b, Bar92a, BPTB76a, Bel94, Ben71a, Ber76a, Ber04a, BW94, Ber86a, Ber93b, Bes94, Bet98, Bet75a, BAA^{+}75, Bet77a, Bet77b, Bet90, Bet65, Bla84, Blo85, Bob65, BFW47, BBF^{+}57, Bou77, Bou61, BBL13, Boy10, Bra09, Bra02, Bri86, Bri87, Bri88, Bro76a, Bro57c, Bro60b, Bro63a, Bro62a, Bro60c, Bro60d, Bro61, Bro63b, Bro74, BM08, BS01, BK03, Bur09, But84, Bye87, Cal60b, Cal84a, Cam79, Car76, Car94b, Car94c, Car82b, Cha58b, Cha75b, CRB^{+}09, Chr71, Cis52, Cle74, CY81, CAH84, C^{+}87a, C^{+}87b, C^{+}89, Coc50, Cof65b, Coh74, CD77, Coh87, Col06, Com74, Com75a, Com57b, Con90b, Coo81, Coo12, Cra83b, Cre70, Cro65b, Dan90b, Dan91, Das68, DPT17, Dav60, Dav79b, Day75h, Day75a, Day77j, DeV62, Dea53b, Dia83, DM24c, DK80, Don85, Don86b, Don89]. Nuclear [Don90, Dow79c, Dow80c, Dow80d, Dow81b, Dow81d, Dow81h, Dow82a, Dow82f, Dow82g, Dow83b, Dow83a, Dow84b, DSZ74, DuB46a, Dud77b, DW90b, Dys60, Eas00, Ede89, Ehr84, Els90, Ema21, Eme71a, Eme76, Eng88, Epp20, Eps87b, Eps97, Eva86, Eva63, Fed58, Far71, Far08, Fed59a, Fel64, Fel67c, Fel74b, Fel76a, Fel83a, FS08, Fer57c, Fes76, Fie87, FM01, Fla87, FL96, Flo76a, Flo76b, Flo78, Fly01a, For83, For87, For09, Foz63b, Fra66, Fra83, Fre84b, Fre09, Fri74a, Fry85, Fry55, Fry65, FWCL20, Gal91, Gal13, Gal56, Gan83a, Gan84, Gen86, Ger84, Gid73, Gil63, Gla76e, Goe79, Gof71, Gol95a, Gol90d, Gol90e, GM09b, GA08, Got87, Got77, Gou47b, GS09, Gra93, Gra84b, GLMS60, Gri85, GC83, Gue80, Gun04, Gus70, Gus83, H.84b, H.85, Hal66a]. Nuclear [Hal66b, HO88, Hal75b, HG84, Ham91c, Ham93, Ham58, Han93, Han98a, Har85b, Hei53, Her85, Her90a, Her82b, Hib95e, Hib95f, Hib06, Hib12, Hin18, HW08, Hoe88, Hof86, Hol74, Hol83b, Hol83c, Hol94e, HK09, Hoo02, Hos87b, Hsi65, Hum83, Hum60, Hum72, Hym07, IK21, Ikl75, Ima70, Ing57b, Ing67, IS67, Ing68b, Ing71, Ing79b, Iri56, Isa92n, Isa02e, Isa03f, Jac71b, Jac72a, Jac73b, Jac74b, Jac80b, JP13, Jas81, Jas91, JM83, Jeh60b, Joh59, Joh60a, Joh83a, Kal64, Kal82c, Kal82f, KHB^{+}05, Kan09, Kas15, Kat54c, Kee76, KS58, Kem09, Ken79, Kha99a, Kha99b, KK02, Khr05, Kib07, Kib12, Kid91, KA90, Kra99, KJE16, Kre07, Kri75, Kri87, KKJK23b, Küh17, Los57, La 88b, Lad81, Lam80, Lan89e, Lap56c, Lap59b, Lap59i, Lap61b, Lap70, Lap73b, Lar65]. Nuclear [Law88, Lay79, Leg56, Len73, Les93d, Lev84b, Lib58, Lib59b, Lif76, Lin84a, LC09, Lon78, Lor01c, Lor02, Lov73, Lym08, M.85, MCAH07, Mac84b, Mak03, ML75, Mar98b, Mas76, McC81, McC09, McD13, McN64, Mec17b, MW77, Mei06, Men48, Met83, Mey86, Mey85, MG08, MR09, Mic62, Mic74b, MW91, Mil85b, Mil12a, Min86, Mol62a, Moo93b, Moo95b, Moo96c, Moo96m, Moo97b, Moo97o, Mor09, Mor85a, Mor84a, Mor85c, Mos92, Mos81b, Mos81a, ML09, ML10, Mur72, Myr75, Nat85c, NR85, Nan71, Nay12, Nei82a, Nel74b, Nel57, Nel92c, Net75, Neu58b, NAD^{+}47, Nit92, Nor55, NA87b, NA87c, NA87e, NA87f, NA87d, NA87g, NA87h, NA87i, NA87j, NA87k, NA88b, NA88c, NA88d, NA88e, NA88f, NA88g, NA88i, NA88h, NA88j, NA88l, NA88m, NA88n, NA88o, NA88k, NA88s, NA88a]. Nuclear [NA88r, NA88t, NA88q, NA88p, NA89b, NA89c, NA89d, NA89e, NA89f, NA89h, NA89g, NA89i, NA89j, NA89a, NA89k, NA89l, NA89m, NA89n, NA89o, NA89q, NA89p, NA89r, NA89s, NA90b, NA90c, NA90d, NA90e, NA90f, NA90g, NA90i, NA90h, NA90j, NA90k, NA90l, NA90m, NA90o, NA90n, NA90p, NA90q, NA90s, NA90r, NA90t, NA90u, NA91b, NA91c, NA91d, NA91e, NA91f, NA91g, NA91h, NA91i, NA91l, NA91j, NA91k, NA91m, NA91n, NA91p, NA91o, NA91q, NA92a, NA92b, NA92c, NA92e, NA92d, NA92f, NA92g, NA92k, NA92i, NA92h, NA92j, NA92l, NA92m, NA93a, NA93b, NA93d, NA93c, NA93e, NA93g, NA93h, NA93k, NA93j, NA93i, NA93l, NA93n, NA93o, NA93m, NA93f, NA93p, NA93r, NA93s, NA93q, Nor94a, NA94a, NA94b, NA94f, NA94e, NA94c, NA94d]. Nuclear [NA94h, NA94g, NBF94, NA95b, NA95c, NA95d, NA95e, NA97e, NA97c, NA99a, NA99b, NA99c, NA99d, NA00a, NA00b, NA00c, NA00e, NA00f, NAKH01a, NAKH01b, NAKH01c, NAKH01d, NAKH01e, NK02, NAKH02a, NAKH02b, NAKH02c, NAKH02d, Nor02a, NAKH02e, NK03a, NKH03b, NK03b, NK03c, NK03d, NK04a, NK04b, NK04c, NK04d, NK04e, NK05a, NK05b, NK05c, NK05d, NK05e, NK05f, NC06, NK06a, NK06b, NK06c, NK06d, NK07a, NK07b, NK07c, NK07d, NK07e, NK07f, NK08a, NK08b, NK08c, NK08d, NK08e, NK09b, NK09c, NK09a, NK09d, NK09e, NK09f, NK10d, NK10e, NK12b, Nye82, Oga58, Oga68, Ore58a, Ore58b, OS00, Osg62, Phy80, Pai82i, Par88b, Pat79e, Pat81b, Pau77, PD89, Per95, Per13, Pet81, Pie73, Pif24, Pit91, Pol69, Pos79, PMS00, Pow04, Pow83, Pre96, Pre62, Pri55, PvH74]. Nuclear [Pri75b, Qua80, Que69, Rab57d, Rab58d, Rab59n, Rab60b, Rab16, Raj10, Ram11, Ram13a, Ram86b, Ram04, Ram05, Rea76, Rei59, Ric09, Ric02, Rin84c, RK89a, Rob91b, Roc81, Roc83a, Rog71, Rom82, Ros85, Ros82, Ros08, Rot58a, Rot78a, Rot99b, Rot02b, Rub08, Rui68, Rus62a, S.55b, Spe58, Sak78, SW07, Say84, Sch74a, Sch74b, Sch08b, STFK09, Sch13b, Sch57b, SS06, Scl79, Sco70, SA62, Sea68, Sea70b, Sha85a, Sha84b, Sha82, Sid00, Sie05c, Sie05f, Sil66, Sim53, Sim54e, Sim59a, Sle84, Smi83, Smi85d, Sne57, SC07, Som83a, Som84c, Spe85a, Spe86, Spe87a, SS90b, STC74, Spi60, Spi65, Spi83, Spo53, Squ12, Squ21a, Bul08q, Bul08r, Bul09a, Bul09d, Ste91b, Ste60a, Ste86a, Ste76a, Ste77b, Ste80a, Ste69e, Sto99, Sto59b, SB23, Sul58b]. Nuclear [Sut95, Tak11, Tay56, TK66, Tel57b, Tel82, Tel91, Ter04, Tob12, Tra90, Tsi83e, Tur17, Van03, Voh70, Wal85a, Wal08, Wal84, Wal81b, War92a, Wat70, Wea79, Web06, Weg96b, Wei53b, Wei54a, Wei70b, Wei78f, Wei82b, Wen53, Wes98, Wes99, WB85b, Wes00b, Wig58, Wig62b, Wil67b, Wil75b, Wil85a, Wil74a, Wil75d, Wil79, Wil81c, Win94, Win98, Woh68, Wol01a, Woo96b, Woo97d, Wou78, Yah69, Yeo75, You65a, You78b, ZC92, Zel76b, Zer13, Zhe90a, Zhe90b, Zin79, Zuc52, vH75a, vH76a, vH76b, vH91, vHZC93, vH00, Abb12, Abr80, Abr84, Abr87, Abt84, Ach10, Ach12, AM86a, Ack88n, Aga79, Agn82a, Akh17, Al 11, AP88, Alb86, AF88, AZ89b, AH91d, AH91e, AH92f, Alb98c, Alf81, AH20, Alv14, ASB^{+}84, APR17, Ano79j, Ano79l, Ano81k]. nuclear [Ano82j, Ano82m, Ano83c, Ano85x, Ano86g, Ano87j, Ano87-35, Ano87-45, Ano87-43, Ano89-28, Ano90-39, Ano90-52, Ano90-58, Ano91-56, Ano92-29, Ano92q, Ano93u, Ano95n, Ano97j, Ano99-29, Ano11d, Ano14a, Ano14d, Ano14g, Ano15a, Ano15c, Ano15e, Arb16, AB18, Ark83d, Ark83c, Ark95e, AG91, Arn96b, AH22, Avi83, Bac81, Bal84c, BW84, BWA85, Bar77c, Bar81b, Bar81h, BW80, Bar92b, BT82, Bel20a, BBP14, Ber81, Bet89a, BH80, Bla11, Bla20, Bol88b, Bor14, BOW10, Boy84, BM22, Bra13b, BMPd89, Bro13, Bud22, Bug22, BBC^{+}98, Can79, Cam90, Car82a, Car83b, Car17, Cav82, Cha85b, Cha89b, Cho81, CN88, Cla77, Cle87b, Coc81b, CAN88, CN90, Coh76d, CEE85, CF87, CM91, Coh96, Col79b, Col79c, CJLT24, Cot19, Dah78, DS47b, Dav79a, Dav82, Day76f, Day77k, DLF79]. nuclear [DM84, Den14, DM24a, Dom82, Dow77, Dow82h, Dow82i, DS12, Doy17, Doy19, Doy22, DG07, Dro16a, Dro16b, Dro23c, DuB62b, Dum76, Eav17a, Eav17c, Edw12, E^{+}84, Eis62, Eps79, ES89, Eps90c, Fre77b, Fal79, Far13, Fed78, FTvHW76b, FUvH85, Fel76h, Fel77n, Fel78k, Fel79a, Fel79n, Fel79h, Fel81b, Fel82f, Fer10, Fie91, Fis86a, Fis81, Fit11, For85, Fos88, Fox88, Fra75, Fra76a, Fra76b, Fra79, Fre83, FC86, Fri82, Fri11, Fur81, Gro78c, GB11, Gan83e, Gap86, GLH^{+}79, GvH23, Gay83, GN10, Gib91a, Gil18, Gla12, Gol90c, GG11a, GG11b, Gra81b, Gre82a, GMV^{+}24, Gri87b, GM14, Gro86, GT14, Gus04, Hen96, Ham98, Han17, HBD17, Han94, HGF17, Har14c, HH87, Hat85, Hau17, HSV19, Hay79, HD23, Hec12, Hec20, HM15, Hec16, Hel11, Her84, Hes84, Hes97]. nuclear [Hic22, HV18, Hol76a, Hol78, Hol80a, Hol16, HK96, Hop84, Hor89, Hos87a, Hug95, HK13, Hun84, Hun90, Int78, Ial18, Ikl76, Ikl80, Ima93b, Ing84, Joh86, Joh80a, Jon13, Jos87, KKL15, Kap85, Kap82, Kaw78, KR11, Kel84, Kel10, Ken87c, Ken63c, Kes89, Khr10, Kib10, KK12, Kim20, Kim22, Kin78a, Kis79a, Kis79b, Kla19, Kom80, Kop15, KR12, Kor17, KSG15, KK19b, KK19d, KK19e, KK19f, KK20d, KN11a, KN11b, KN11c, KN11d, KN11e, KN12a, KN12b, KN12c, KN12d, KN13a, KN13b, KN13c, KN13d, KN14a, KN14b, KN14c, KN14d, KN14e, KN14f, KN15a, KN15b, KN15c, KN15d, KN15e, KN16a, KN16b, KN16c, KN16d, KN17a, KN17b, KN17c, KN17d, KN17e, KN18a, KK18b, KN18b, KND18, KN18c, KN18d]. nuclear [KK19c, KK20a, KK20b, KK20c, KK21a, KK21b, KK21c, KK21d, KK21e, KK21f, KK22b, KK22a, KK22c, KK22d, KK22e, KK22f, KKR23a, KKJK23a, KKJ23a, KKJ23b, KKR23b, KKJK24a, KKJK24b, KKJK24c, KKJK24d, KKJK24e, KKJK25, Kro15, Kru81, KSS17, Kul17, La 79, LR79, Lan16, Lan82a, Lan90j, Lap76, Law85, Law11, LL83, Lee24, Ler81, Lev12, LL12, Lif80, Lif85, Loc14, Lod84, Lon76a, Lon84, Lon85a, Lor77, Lov86, Lov13, LJ17, Lun82, Lym11, Mac11, Mac81b, Mag16, Mal85, MFN17, Man77, Man84a, Man84c, Mar76, MB98, MDB81, MK02a, Mat83b, Mat12, May17, MR77, McM82, Mec12, Mec13a, Mec13b, Mec16e, Mec16f, Mec19e, Mec19f, Mec20a, Mec20b, Mec20g, Mec20h, Meh21, Mei20, Mey83, MN96, Mia15a, Mia16, Mia15b, Mic81]. nuclear [Mil79, Mil86a, MM90, Mil10, Mil87, MW98, Mir89, Mit58, Møl90, iM16, Moo96n, Mor82b, Mor83c, Mor84c, Mos80, Mou17, Mus82, ML82, NPS90, NP80, Nel81, Nel10, Nel12, NF83, NH08b, Nol17, NCA85, NA87a, NA87l, NAC87, NCA88a, NA92p, NA95a, NA95f, NA96f, NA96d, NA96c, NA96e, NA96b, NA96a, NA97a, NA97d, NA97f, NA97b, NA98g, NA98d, NA98e, NA98f, NA98b, NA98c, NK10a, NK10b, NK10c, NK11, NK12a, NK13, NK15, NK16, Nye77, Nye13, O’B79, O’N89a, Oel14, OW15, Pon82, Pai83c, Pan23, PPE22, PE24, Par78b, Par17, PFW18, Pat78c, Pat79b, Pat80d, Pat90a, Pei82, Pei84, Per91b, PP83, Per20, PD12, Pif18, Pin24, PR85, PB^{+}01, Pod16, Pod24, Pol80, PN09, Pon17, Pop93, Pos85, Pra89c, R.83, Roy76, Raj17, Raj14]. nuclear [Ram13b, RM16, RK19, Rem15, Rep10, Rin84d, Rob89, RT12, RTB^{+}19, Roc76, Roc18, Rog88a, RS16, Ros89a, Ros87b, RG13, RL14, RH17, RE19, RF21, RH12, Rot77, Rot93o, Rot94d, RAB^{+}80, Saa84, Sai10, Sal79a, San76b, Sar17, Sas17, Sau10, Sau17, Sch15a, SFT11, SF12, SF14, Sch78b, Sch98, SW59, Sha87b, Sha18, Sha85c, Sha93b, She94, She16, She14, She85c, Shk92, Sho80, Sho92a, Shu81, Sie80, SW21, SL80, SD15, SOE21, SS76, Smi85a, Smi90b, Smi95a, SM14, Sno81, Sok16, SR81b, Sol90, Som84d, Spe89b, Spe78, Squ13, Squ16, Squ17a, Squ17b, Bul09h, Ste89b, Ste81a, SHT^{+}87, Ste77c, Sto16, Sto17b, ST22, Sub77, Sub84, Sub87, Suz11, Swe77, Swe79a, The89a, Tar23, TBBS85, Tay86, TD89, Tho80a, Tho81a, TA94, Tho86b, TY75]. nuclear [TT83, TPPB21, Tre19, Tsi82d, Tsi83c, Tur77, Tur18, Tur13, Uni78, Ust93e, Utr87, Vad21, Wag88, WH89, WL84, WH11, Wal21a, Wan83, WSG^{+}15, Wea92, WP19, WW60, Wei76a, Wei77a, Wei80a, Wei86a, Wei78b, Wei91b, Wei15, Wei84, Wel80, Wel16, Wes79a, Wie84, WM78, WF90, Wil77c, Wil80d, Wil82c, Wil86b, WF16, Wol80, Wol91a, Wol01b, Wol17, Woo86, Yan82, Yin15, Yor76a, Yor81, Yor82, Yor83, ZGD83, Zal19, Zei85, Zha12, Zha15b, Zha19, Zik21, ZG83, Zin53, Zuc82, Zwe84, Zwe85, Zys17, vHW76a, vH79, vH^{+}90b, vH16, And65, Ano60f, Ano75g, Ano83c, Ano83g, Ano88e, Ano91s, Bar80g, Bas80, Bre93, BHK^{+}21, De 70b, Dow82h, Dow83d, Fel68c, Fra83, Gan83b, Hol66, Hoo20, Hos87b, Kal82b, Kap74, Mor62c, NA87d, Pat81a, Que67, Rob20, Sch63a]. Nuclear [She61, Smi83, Ste76a, Sto19c, TB23, Wei78e, MD81, Abb85, Ack87a, Ano57d, Ano84g, Ano85b, Ano85c, Ano87o, Ano89f, Ano98h, Bar73a, Bar80m, BD79, Ble87, Bre93, Cal84a, Cha83b, Dav79b, DeN74a, Don90, Dow77, Dow79b, Dow82f, Dow84d, Eng88, Fei77, Gan83b, Gan83c, Gil83, Gol83c, Han94, Hei54, Hol77, Hos87a, Lap56a, Law88, Lin85, Mar78a, Mar54f, Moo97a, Moo98b, Mos81b, Nat87, Par80, Par86, Pol68, Pre96, Rub73, Sac84, Sha84b, Smi88, SB91, dF79b]. nuclear-age [Lif85, Rin84d]. nuclear-armed [Cot19, McD13, WF16]. Nuclear-capable [NA87b, NA87l, NA88p]. Nuclear-Free [Das68, Don85, SB23, Ano89-28, Ano90-39, Fry85, Gri85, NA93e, TD89, Zei85]. Nuclear-Power [Abr73]. Nuclear-Powered [Ano55g]. Nuclear-Reactor [Pri75b]. Nuclear-Renewable [For09]. Nuclear-Weapon-Free [Moo93b, Bul09a, Wal08, ML82, She94]. nuclear-weapons-free [And85]. nuclear-zero [Rep10]. Nuclearisation [Jon90b]. Nuclearism [Keu83]. Nuclearizing [Eis59]. Nucleonic [Ano49g]. nudge [Coh93b]. Nuell [Wil69b]. nuff [Ano93v]. Nugget [Wel79]. Nuke [Sch01b, Ano98n, Ark94b, Cle92, NA94f, Wil15, Ano94m, NA91l]. NUKEMAP [Eav17c]. Nukes [Ahm00a, Ark00b, BC96, HT03, NA87h, NA92n, Opp03a, Opp04, Sch02a, But97a, Coh93b, FM93, IS92, KK98b, NA92h, San14, You92]. nukespeak [Lan90f]. Nukey [WM78]. Nuking [Ark96c]. Nukleare [Ano75g]. Number [Ano75d, Cat87, Chu94b, McC86b, Opp02b, Rot98j, Sch75e]. Numbers [Ark02b, Asp80, Bal79b, Bro85e, Cat89, EB73a, Gou78a, NA89p, Pol73, Sch95a, Ben83, BL83a, Bro90, CB79, Day76g, Eri88a, Eri88b, Lei82b, Loe88, NA90j, Rot98i, The89c]. NUMEC [GM10]. Numerical [BCW74]. Nunn [CLD88, Isa89c, Isa90n, Isa91q, Kre88e, Loc93b, Loc95, Pai85b, Rob91a]. Nunn-sense [Rob91a]. Nuplex [Mei69, Sha69]. Nursery [SG65]. nurtured [Dav77a]. Nurturing [AB08]. NUT [Cou81, Lin82]. Nutrition [Spu73]. nuts [O’S86, Rot98a]. Nutting [Ark87c]. Nutty [Ano92q]. Nuuk [Spr19]. NWFZ [HW96, Mti96a]. NYC [SB23]. NYCAI [Lon50]. Nye [Cra84c, Nat87, Wal85a]. nyet [And93, Ano91-56, KOE93, Dav92].

O [Ano78m, Bro47c, Deu83b, Dob54, Fel81a, Lap69, She60, Smi58f, Sol68, Sta85, Tho80g, Wun82, Sin14, CO03]. O.K. [Ano96m, Rot92n]. Oak [Ano46-69, Ano48t, Ano49-33, Doy17, McE52, Nor47, Ric74, Whi48b]. Oakland [NA90k]. oath [Lem88]. Obama [Mec24, Bul09a]. Oberdorfer [Cum98]. Obey [Fah91]. Obituary [Ano46-70, McM03, Rab52g]. Object [Bul01b, DSL^{+}87]. Objecting [Ano49-35]. objections [Fin77b]. objective [Sie82]. Objectives [Bla87a, DuB63]. objectivity [Bla77]. Objects [Sag67, Hyn69]. Obligation [Dre74, Ein52a, Osb54, Ein79b]. Obligations [Ede58, Mas48b]. Oblivion [Ano74c, Lei74, NB77, Dow83f]. O’Brien [Bri86]. Obscene [LM75]. Observation [Lal65c, Pom67]. Observations [Dub55a, Lof60, Shi57, Wil66]. obsfuscations [Fel79f]. Obsolescence [Boy89b]. Obsolete [Bot69, Mac51, Mon54, Fel80j, Wei84]. obstacle [Ema21, She86a, Wie83]. Obstacles [Hei11, MCAH07, Wei64a, Dav77a]. Obtain [Pau52b]. Obvious [Fre78]. Occasion [Aur71a]. Occupational [Pra05d]. Occurrence [Bra50, Doy17]. Ocean [Bal79c, Che70, Lei70, McK74, Pon67a, Ste73b, Bar80f, McK74, Woo73b]. Oceanic [Bre74, RE68]. Oceanography [Spi64, Chr68]. Oceans [Dra73, Cot19, McK74]. O’Connell [She89]. October [NA91a, NA91c, NA98e, Rab56h, Rab70c, Ano56m, Ano57j, FR76, Gui65, Ing65, Lal65a, NA88b, NK12a, Sch89c, Zag93b]. Oda [McK74]. Odd [Isa88c, RK88, Fla90]. odds [Kre87b]. Ode [Fel77h]. odor [Kre91a]. Odyssey [Isa05a, Woo98e]. odysseys [All16]. OELs [Bal87]. Off [AZL90b, Dil57, Hib01, Isa03f, Mae01, Myr75, NA88b, NA88k, Pri07, Wal06, Woo00f, Ano89k, Ano94o, Doy22, Gad89, Mak90, Per90, Rot89b, Rot93r, Rot95g, Sez16]. off-ramp [Doy22]. Off-Shore [Woo00f]. Off-the-Shelf [Pri07]. Offense [Rab66c, Abl85, Ano92-30, Kre92a, Sen87]. Offensive [Gar67, Sei53, Nei88, NAC87, Cla54]. Offer [Bur50, Ano98o, Lea89]. Offered [Gal63]. Offers [Kih47f, Ano87-30, Sca88]. Office [Ano45c, Ano47i, Ano48z, Moo05a, Wel08, Ano46-79, WC47]. Officer [Gou55, Lef77a, Mec16b, Mec24]. Officers [Wei85b]. Offices [Wei85b]. Official [Ano62u, Ano64p, Ano85o, Cha72, DMD64, Gor51b, Mer46a, Mil48a, Ano97j, Bat86, Gaf86, Kra92, Pat86b, Tim87]. Offner [Bou02]. offset [KE17, Squ17a]. Offshore [Dro21c, Ing79a]. often [Jam86]. Ogburn [Pla57b]. Oh [Ano92u, Ano92q, Ano94n, Pri03, RL96, Rot94f, Ano96o, Aue02]. Oh-So-Late [Pri03]. Ohio [Aue02, Lor01b]. Ohlson [Ano86i]. Oil [Ano08h, Ber77a, Cav04, Cav05, Coo75a, Por74, Pre80, Sin05a, Ste86c, Wou81, Ber78b, BC16, Cav13, Isa94i, Red82, Sez20, SCHD20, Wil80c, Ust93a, Fla80b]. OK [Isa93d]. Okazaki [Dow82g]. O’Keefe [Gol83c]. Ola [Dav79b]. Old [Ano79l, Ano07p, Bla61a, Cal61b, Ded70, Dos72, Gou63, Han01, Kal61, KD59, Kel87, Mel00, Nin88, Ric86a, SS69, She85e, Web03, Weg96a, Afh88, Ano93z, Ano96z, Ano14c, Bro90, Hof88, Kre91a, Kre91b, Lon87, RD91, Wit95a, Zag95a, Zag95e, Zam92b, GC63, JvL63, Lar63, vL63]. Old-fashioned [Kel87]. oldies [Ano98o]. O’Leary [Gra94, Sal71, Wee98]. Oleg [Ano90o]. Olga [Ekl80]. oligarchs [Har22]. Olive [Ano87u, Ano92l]. Oliver [All55a, MJ68]. Olli [Ano11d]. Olof [Ano86-30, Tra72a, Wei86e]. Omaha [Ark00d]. Ominous [Day75n, Ark84e]. Omission [Bro47e, Den47, Red74, Ris75, The89e]. Omni [Ano85d]. on-site [GK87]. Once [Alb64, Ano67e, Ano69e, Moo93a, Wei78g, Ano95n, Bro85i, Loc89b, Men97].

One [Ama77, Ano55m, Ano80q, Ano90-42, Ano96o, Ano98o, Ano13b, Ark99d, BPR52, Ben72, Ber01, Ber72d, Bri86, Cai89, Cha75d, Cir05b, Cle04a, Cle04b, Cot22, Dje74, Eva84, Fin53, Fri20, Gol90b, Har69a, Hit03, HE72, Hul76, Isa89i, Isa02b, Key66, Kri02a, Lan90g, Lap51b, Lor70, Man64, Mat80a, Moo96h, Mur55a, Pai04f, Pal01, Rot02c, Ryb71, SW80a, Tur51, Wad74, Aba89, Ano88v, Ano91-35, Ano94n, Ano95j, Ano95n, Ano95m, Ano98m, Cat87, Coh98, Col79b, Eps93, Fed76, Fel80d, Gar97b, Har10, Hil16, Hil20, Isa88e, Isa89l, Kis76a, Lym11, Pen98, Rof19, Rot18, Rot98j, Spe95a, Wur93d, Zal19, Zam98, Zim81a, vV23, vH80, Dow83f, Ano46u, Lew72b, WM82, Win94]. One-Dimensional [HE72]. One-half [Eva84]. one-liners [Isa88e]. One-Man [Har69a, Hil16, Hil20, Spe95a]. One-Sidedness [BPR52]. One-third [Dje74]. One-time [Ano98o]. one-two [Lym11]. One-Vote [Isa02b]. O’Neill [Boy95]. Ones [Isa01c]. Ongoing [All06]. online [BV21, YWS20]. Only [Ben72, Dal75, Dav77c, Hur75, Rab51g, RR07, Rus58b, Sto56a, Ano14g, Cla77, Coc89, Fly98c, KS94, Kas77, Kla19, Nel94, Pie22, Rot93i, Ste77c, Wat77, Ber72d, Rus20, Kos60]. Onofre [Sch74b]. Ontology [Mar49b]. Onward [Isa00a, RWK^{+}83, NA93j]. Oops [Ano98k, Han90c, Han00b]. OPEC [Cav13]. Open [Ad 90, Ack89f, ADH^{+}73, Ano62v, Ano62w, Ano67j, AM74, Boh50, BH65, Cre90, Ein49, Han95c, Hay55, Lev64, LPT^{+}54, Mak95b, Mar02b, Opp49b, Phy80, Rab62c, Rab62j, Rab66b, Sak75, Sch63b, Spe89a, Str88, VFIS48, Ano91-41, Ano94m, Ano95l, Dan91, Gaf86, Gol90e, Gol91c, JH16, Lin88, Nat81, Nat85b, NA90k, BBG^{+}85, Pat80e, Smi92c, Dia84]. Opening [Ano64n]. Openness [Ste81a, Zai06, Ano95k, Ano95m]. Opens [Foz61i, Kih47f]. operating [Bra17]. Operation [Ano49e, Rab53f, Rot95e, Chi69, Con46b, Wei13, Ano46-71, Das96, Fuc70, NA92f, Owe54, Pfe82, SH94]. Operational [NA94c, dRVT^{+}47, Ram16]. Operations [Ano53t, Ano88c, Bos51, Kat52a, Mec16b, Nat86a, McE52, For87, Hul98]. operatives [Kla88]. operators [Hor89]. Opinion [Fel67b, Haa73, Joh97c, Kra72, Lau46, Mur54a, Pom61, She97, Smy54, Ste75b, Str54b, Ano76d, Ano98l, Arn97a, Ber95, Bla97, Bun98, Cra96, Cum94, Gey94, Her95a, Jac82a, Kre97, Lea97, Mac96, Mak95b, Men95, Mey94, Mot82, Pap98, Pik96, Ple96, PY98, SW56, Sha96, Sim95, Smi95a, Smi98, Tay96, Wei94c, Zam98, Alc63, Ano81c, Val86]. Opinions [Ano46-81, BB52, Ein54, Kih47f, Lov74b, Min48, Bye88, Som20]. Oppenheimer [Bet67, Gro56b, RB67, Wea82, Zac76, Ato54a, ACU^{+}54, Ano52r, Ano54v, Ano54w, Ano54z, Ano54-28, Ber04c, BS05, Edi54, Bro23, CH54b, Coo04, Dav68, Dav86b, Dro23l, Dro23k, Exe54, Eva54, GNMS54, Gar54, Gol92b, Gol55, Gre77, Gre23, GDH^{+}04, Haw94, Kal54, Kal63b, Kam23, Lev60, Lif23, Lip54, Mac54a, McM05, Mec23a, Nic54b, Opp51b, Opp54b, Opp79b, Opp23, Sim54b, SW80a, SW80b, SG69a, SG71, Uni54, WS79, Yor76b, Yor89, Wil69b]. Oppie [Dro23b, Sch04b]. Opponent [Has68]. Opportunism [Wri49]. Opportunities [Ano55j, Bro55c, Sch00b, TPPB21, Ano92z, Gro15, Sha95b, Ano78m, Ore58a, Chr68]. Opportunity [Ale64, Aus47, Isa05d, Wol75, Ack87b, Car84b, Fal87, Kri97, Mac86c, Sez16, Sub93, SC97]. Oppose [Ano51z, Lap76]. Opposed [Joh83a]. Opposes [Ken60a]. Opposing [Ber49, Coh74, Pol49]. Opposites [Cah94]. Opposition [Bou77, Maw76, Sig87b, Swe77, Zam98]. Oppression [Ote81, HM15, OS81]. Oppressions [Mey69]. O’Prey [Uhl96]. Optimism [Abb93, Pre63, Ree07, Rus55b, Woo78, Fig23, Hel11]. optimist [Ano10b]. Optimum [ILA70, Woo81]. Option [Ano89g, Ber71, Car94c, Fri74a, Rex71, Yah69, KS94, Kün96, Moo93b, Pan98a, Ram10, Dow83c, War92a]. Options [Ano88p, Maw77, Moo96i, Flo78, Gri85, Hei11, Mak97, Nat81, Red78a, Red78b, RL78b]. Oran [Jas81, Lal67a]. Orbit [Ber73a, Bro89b, Woo01c, JR89]. Orbiting [Smy64]. orchestra [Zag93a]. Order [Ano55w, Ano75c, Elm66, Fel84b, Hur78a, Jac72b, Kas64, Lan61a, Pag68, You67b, dF79b, Aja78, Ano91-67, Ark95d, Bud22, Dea91, Dow91, Fah91, Han17, Hui91, Ive16, Joh91, Kap91b, Kre91a, Kul17, Men91a, NK12a, Per91b, Pol91a, Raj17, RK91, Sok91, Sol91, Ste91c, Woo91, You91, Ano78n, Wil84, Hoe88]. Orders [Ano50-37, Liv68, Ano94s, BT94, NA91b]. Ordinary [Hil51, Dro22g, Dro23b]. Oregon [La 88c]. Ores [Ano48n, MEA^{+}47]. Organic [Dob64]. Organism [Dob54]. Organization [Ano45h, Ano45m, Ano47-27, Ano49-39, Ano50-36, Ber66a, Dew55, Ehr82a, Gre73b, Gro50, Hir55, Lil47c, Pow47, Rob66, Sho46, Spi72a, CAM20, GK21, Ano47-27, Noy46b, Zer13, Hur78a]. Organizational [Cha69b]. Organizations [GS51, Gra84b, Wei53a, Yar80a, dF79e, Bou85]. Organize [Ano49-35, Shi47d]. organized [Cro17d]. Organizer [Her02]. Organizing [Dum93, Gar66, Gar60, Jac63b, Sch11, Mor62d, Sto59b]. Orient [Kom54]. Oriental [Lal75]. oriented [Pan23]. Origin [Fox73, O’K73, Ure69a, Ure69b, Ure73, Wes00b, Dob54]. Original [Sie05f, Thi46]. Originalliteratur [Thi46]. Origins [Boy87, Dar86, Fel73d, Fel77o, Mei51d, Pat87, Pre75, PvH74, Wea79, Yor75a, Har86a, Her87, Gra62]. Orlov [Ano78q, Mar78b]. Orphan [Joh97c, Ano90-37]. Orphans [Ren99]. Orr [Jac73g, Smi85d, Ano51-29]. Orthodox [Woo98e]. Orwell [Ano86z, Fel83b, Nat84, Rot92g]. Osamu [Ano13f]. Osborn [Ano47w, Gro47b, Pog54]. Oscar [Nie63]. Osgood [Roc81]. Osiatyński [Ran85]. Osirak [Fai81, Smi81, WK82]. Oskar [She60]. Oslo [Ore61b]. Ossietzky [Sta79]. Osterling [Kat52b]. OTA [Ano88g, Cas78a]. Other [AH91l, Alf74, Alv74, Ano47u, Ano69l, Ano69k, Ano78o, Ano78m, Bor52b, Ded66, Dys63b, Fel69b, Fer52, Han69, Hel07, Kis69, Lep89, Mer79, Min52, Pow74, Rab69a, Sie06d, Smi83, Uni57, dF79h, Alb98a, Ano90-30, Ano98m, Cla90, Day77d, Eds76, Gor95e, Har85a, Har86b, Har86c, Har87a, Har87b, Har88a, Hor88c, Ler80a, Mil95, Mit83, MJD92, Pai85d, Kün96]. otherwise [Pen98]. Ottawa [Lea97]. Otto [Cro65b, Fel79o, For80, Kap61, Hah67a, Hah67b]. Ought [Ber49, Pol49]. Oughterson [Has57]. Oui [FC96]. Ounce [LH05, Ano96n]. Our [All16, AC49, Amr54, Ano46f, Ano53-36, Ano53-44, Ano63v, Ano72i, Aus47, Bai75b, Bai75a, Bro54, Cha68b, Cle71, Day75g, Day77l, Dow82e, Dow84d, Dow84c, Fer57a, Fit75, Gou48, Hal75a, Hel07, Kam74, Kin67, Kis93a, Ler53, Lew69b, Lin55, Liv51, Mac54b, Man74, McD74a, McM48b, McM49b, McM52, Mea62, Mey18, Mul72, Olt52, Opp47a, Opp56, Oro59, Osb54, Osg67, Pec70, Pla78, Pow90a, RG49, Rab62e, Rab69i, Ros52, Rus52b, Sch59b, Sea68, Sea70b, She61, Sim57a, Wat74, Wei46, Wel60, Whi53, Why71, Wil74a, Wil75c, Zac50, Zif53, dH75, And74b, And74a, Ano97n, Ark89a, Bac77a, Bai20, Bas80, Bel20a, Che20, Dro22b, Etz77, Gre83a, Isa93o, Mec20d, Mil89a, NA91h, Nye13, Opp20, Qua80, Ran82, RW76, Rot96f, Sal80c, Sca11, Sob21, Soc15, Tha63]. our [Wei17, Wil75f, Wil75d, vH79, DuB63, Lan90l, Tod64a, Ano46v, Ore58a]. Ours [Bot61, Lif23]. Ourselves [Hen93, McK07, Why71, vH00, vW58, Dah77]. outbreak [SAB^{+}16]. outen} [Ano93v]. Outer [Ano58j, Bal79c, Blo63, Bre59, Che70, Cle67, Gal65, Gor67a, Lal66c, OR70, Pre58, Sch62a, Tho80d, Arb19, Arb85, Kor19, Lay79, JFB19, Deu83b]. outgunned [Sch00d]. Outlaw [Her95a]. Outlawing [Sho92a]. Outlaws [Moo95b]. Outline [Kam71]. Outlines [Bra49]. Outlook [Bir85, Bro49, Cra85, Eme71b, Fed72, Isa85a, Kla85, Pea85, Rab57g, Rab63f, Sea73, Spe85a]. Outlooks [Bro57c]. Outmaneuvered [Sch00d]. Outmoded [Ich64]. Outrage [Pal72]. outrages [Isa93k]. outs [NA91l]. outsider [Sto19c]. Outsiders [Mey05]. outspends [Hay81]. Outspoken [All60g]. Outstanding [Bal89a]. outweigh [DS47b]. Outwitting [Got69]. Over-Emphasis [Kep76]. Overcome [Sch54c, Con90a, NH08b]. Overcoming [Ano78d]. Overdue [Fly01b]. Overholt [Jac80b]. Overkill [Blu84a, Gre63a, Kal63a, Khr63, Lap63, WMT93, Day77b]. overlooked [Jam86, Pet11]. overprogramming [Gor95f]. overrated [Fla88, Kam96]. Overreaching [Sca91]. Overseas [San66a]. Oversell [Isa90t]. Overshoot [Ehr81]. oversight [Ano89s, Ano96n]. oversold [Lon76b]. overspending [Mec21f]. overview [Ack88n, CAN88, FS84, Gol24, HK18, La 83, RE19]. Overweight [Pri03]. overwhelming [Hol20, Wei80c]. Ovshinsky [Ano11h]. Owen [Gri83a]. Owens [Ber78a]. Owls [Wal85a]. Own [Ama03, App53, BCMR92, Dow82e, Lor70, Nea02, Pow01, Rab53f, Ryb71, Tuc08, Ano97n, CKN^{+}94, Coo78, Gra81a, Mei20, Zag93c]. Own-Again [Ama03]. Oxford [Nor06, Jac63b]. Oxygen [Vol74]. ozone [Bru15, Dro23e]. Ozzie [Ano12e].

P [Ano75e, Ano78k, Ano78l, Ano85c, Ano86j, Ano13a, Ben07, Bro57b, Cav57, Cop65, Das96, Eck78, Fal61a, Fel81a, Fla68, For80, Gan83b, Gil85, Gro56b, Gus83, Han75, Hol86a, Lan55, Mar72a, Mar94, Mil85b, Mor62d, Pal64b, Pan63, Per95, Pla59a, Sti83, Tho80g, Uhl96, Cra67a, McD13]. P-M [Cra67a]. Pace [Ben58, Fie21a, Spe85a]. Pacem [Ber84b, Pop64]. Pacey [Pow91b]. Pacific [Ham91c, Ark83c, Ark87h, CS17, Dan90b, Fry85, Gus69, Joh80a, LC78, Mac54a, Mar68, MR77, Moo97j, NA88l, NA89m, Str54b, Sut04, Tam74, Ols62b]. pacification [LD89]. Pacificism [Eva75]. pacifism [Kun97]. Pacifist [Pas55a, Ano88u, Rot79b]. Pack [Lum99]. Package [Isa91j, Ano87-44]. Packing [Rab53c]. Pact [Jac74b, Sta85, Moo96l, Ano87-31, Fis83, Sha85b]. Pacts [Van46, Man89]. padre [Ano92u, Axt73]. Page [Ano53q, AH91m]. pages [Ack97, Ano92-27, Ano92-34, Nor06, Rot97c, Tuc97, Wil71, Bou02, Wil71]. paid [Gad89, Rot89b, Win79]. Pain [LC98b, vD78]. Paine [Wal81a]. Painful [Wad74]. Painless [Lew70c]. Pains [Pra06b]. Paint [Led94, Pri02]. Painter [McN59]. Painting [Lan90h]. Pak [Per95]. Pakistan [Tar23, Ahm00b, Alb87c, AZ89b, AH92i, Alb98a, Atr06, Bla11, Gau24, Haq05, Hus04, Kha80, Kha98, Kha99b, KN11c, KKJ23b, Mus00, Mus01, Mus02, NA98a, NA98b, NAKH02c, NK07c, Rao87, Reh14, RTB^{+}19, Sal64a, San14, Sas17, Wei04, Wil82c]. Pakistani [Ano07q, KN15c, KN16b, KND18, KK21c, Mal14, NK09a, Spe87b]. Palaeoclimatology [Ada61]. Palau [Sca91, Sca88, Sca91]. Pale [Ano99d]. Palevsky [Moo00a]. Palisades [Gen71]. Palme [Ano82p, Ano86-30, Tra72a, Wei86e]. Palmer [Kat54f]. Palomares [Mor09]. pals [Ano90-32, Pea90, Zim89b]. Pam [Swe89a]. Pamela [Ano85d, Dow82e, Dro20b]. Pamphlets [Ano59h, Ano60s, Ano61u]. pan [Wea92]. Panacea [Cra84b]. Panama [Hua02, Lew68f, Pug02, Woo97e]. pandemic [Cro18, Sch20, SCHD20]. Pandemics [Smi21]. panderers [AH92j]. Pandora [Ise92]. Panel [McE52, Bet86, Ano56j, Fal79, Fra79, Ken79, Kis79b, LR79, Sal79a]. panelist [DvH91]. Panic [Ano54a, Hue04, Mar54c, Mer54b, Wil72b, Wyl54, Ano94m, Dal22, Fel77g, Hib94d, Moo96b]. Panjandrums [Mac68b]. Panmunjom [Kap91b]. Panofsky [dF79a, Pan00]. panopticon [Dro22f]. Panos [Her90b]. Pantex [NA92a, NA92h, NA92n, NA93s]. Pantexes [NA93f]. papa [NA91b, NA91m]. Papal [Pop52, Fox88]. Paper [ACH84, Ing52a, NA88h, San57, Shi52, Shi87, Tod79, Ack97, Boo22, CK20, Isa93j, Rab60h, Sie05g, vMC98]. Paperclip [Kor92, Sea72b]. Papers [Goo00, Rab72a, Ano84k, Tod78a, Rot99b, Cha74, Gab73]. Papp [Sha91b]. Papua [Sut01]. parachutes [Pow90b]. parade [NA89j, TM96a]. Paradigm [Dun83, Nag81]. Paradigms [Kol75]. Paradis [Dro22f]. Paradise [Joh80b, MJ68, Moo96c, Sca91, Sch90, Sut95, Bar81b, Alc81, MJ81, RGA81]. Paradox [Gou75b, KMN04, Ram06, All15b, Isa85a, Luc90b, MM86, Mow95a]. Paradoxes [dF79g]. Parallel [For85, Har50a]. Parallelism [For80]. Parallels [SJ68]. Paralyzed [Pet11]. Paramount [Ure49b]. Paranoia [Den77, Sly99]. Paranoids [Tob77]. parasitic [Hum77]. Parceling [Loc93b]. Parenthood [DIJ73, SB73]. Paring [vH97b]. Paris [Gom70, Mil81, Noy47, Pan81, Que70b, Whi48a, dR49]. Parish [Hol94d]. Parity [La 78, Szi64, Szi65, Khr96a, Nit89, Fri59]. Park [Ano46-90, Ano94k, Ano96m]. Parker [Par73]. Parkinson [Par62]. Parliament [Ali57]. Parliamentary [Ano58c, Orr56a, Pri53]. Parliaments [Ded78]. parrot [Kre91b]. Part [Ano48-28, Ano48-29, Ano79m, Ano80q, Ano91-40, Bou00, Cra70a, Dan70, Fuc70, GJ84, Hah67a, Hah67b, Ima70, Isa02a, Kri02a, Lov74c, Mor84c, Pot70, Rom83, Sea70b, SW80a, SW80b, Tsi83d, Tsi83e, Wei70b, Yar80c, Zub00, Ano90-34, Ano91-57, Ano95m, Ano97k, Ano97l, Ano97m, Ano98m, Eze16, Gra16, Han95a, Tal18, Wan16, Adl57, Ano46-51, Ano46-52, Ano49c, Ano53-43, Ano54-31, Ano69o, Ano69n, Ano69p, Ano69m, Ano71c, Ano73f, Ano74h, Ano79g, Aug56b, Bat46a, Bat46b, Bel57, Bet70, Bos51, Bro56c, Coo56, DB52a, DB52b, DB52c, Dan54b, Dub56b, Ell56, Fin56, For51a, For51b, Fra57, Fri56, Gal56, Gom75a, Gom75b, Gro56a, Had56, Hol83c, How56, Ing56a, Kal60, Kal61, Kap78c, Leg56, Lew56, LC78, Lon52c, Mar83b, Mey56, Mor57b, Mor83c, Nea78b, Oom46a, Oom46b, Orr56b]. Part [Par46b, Par47, Pol57b, Rab56g, Rab56k, Rab56l, Rab61g, Sar83, Sha78b, Sin56, Ski53, Tsi83i, Tur51, Vis56, Wei56a, Wid61a, Wid61b]. Parted [Dav79b]. Parthasarathi [Ano78l]. Partial [MW47, SL87]. participant [Pei85]. Participate [Spo51]. Participation [Ano85d, Lau47, Wol68, Yar80a]. Participatory [LEH^{+}71]. Particle [Ano57j, Gar78b, GR79]. Particles [Ano78m]. Parties [Ame69, ZS93]. Partisan [Bin54, Hed65]. Partition [Mul65, Sch91, Van91]. Partly [BB63]. Partly-Baked [BB63]. Partner [RK89a, Wit01b]. Partners [Isa90o, Yaf06]. Partnership [HB10, Hin18, Mul65]. Parts [Goo05, Ric01a, Dow81d]. Party [Adv60b, Ano48h, Ano60o, Ano60n, Ano60p, Ano60q, Bal88, Fre60, Hol86c, Kal70, La 66, Pai04d, Rab59k, Woo99b, GS96, Ken87c, Lon92, NK20, Rot92o, Gap86, Tho80e]. Pascal [Ano98g]. pass [Ano94n]. Passage [Fel86a]. passé [Ano91-55]. Passed [Ano46-83, Ano50-43, Mil48b]. Passes [Ano46-55, Fel76c]. Passings [Ano05p, Ano06p, Bul08p]. Passion [Pol57b]. Passive [Hoh75, RLG11]. passivity [Eva75]. Passport [Che56, Pau52b]. Past [Ano90-34, Ben01c, DeW62, Hir72, Lap70, LvH18, Lov00, MD64, McK07, Pal55, Rab62d, Ric98a, Squ77, Van46, AG97, Ano97q, Bro85d, Cra82a, Dou24, Hol04, Mcc88c, OW15, Puz93a, Rof97, Rot97b, SM14, Tsi82a, Sto55a]. pastoral [Mil83b]. Pasztor [Rot96b]. Pat [Hur78a, Coh87]. Patch [Wea99, Nel77]. Patent [Ano48b, Ano63z, Ano71h, Lis46b, Mil48a, Oom46a, Oom46b, Wel08, Wri63b]. patented [Ano95l]. Patents [Elm65, Hol64, SL46, Wri64, Les81]. Path [Cob78, Dou68a, Dre07, For09, Lew70c, Rob91b, Ber85, Cal24, Cas77b, Ege16, Ike16, Kas15, KR11, Lan16, Mac86b, iM16, Moo92j, Pea91, She16, Sid16, Dow83c]. pathogens [BG14, Esp14, Hun14, Tom14]. Pathology [McA76]. Paths [Kli81, Nef78, AH91l, Eps93, Fel14]. patience [Kim95]. Patient [Got69, Lev77b]. Patricia [Ano78n, Tod81]. Patrick [Dow80g, Ise92, Men91d]. patriotism [Ano91-42, Ber95]. Patriots [Har88c]. Patrol [Kou01, Isa90o]. Patron [Can70]. Pattern [Ano60f, Bat46a, Bat46b, Gen71, McE70, Stu69a, Har84b]. Patterns [Gol95d, Mor62d, Sch70a, Shi56b]. Patterson [Ano46-91, Mil85a]. Paul [And78a, Ano85d, Ano87o, Ano87n, Ano89e, AM86b, Cah94, Cor94, Dev69, Don86b, Dow80h, Eck78, Gor73, Har56, Har81, Hos87a, Jor87, Kre90, Moo02d, Mor82a, Mor84a, Mor62d, Nei82a, Nor02a, Sch86, Sha90, Smi85d, Swe88, Tho62, Tir88, Tod80, Tod81, Tra71, Wol63, Ano11e, Mec23c, Sto17c, Dow79c]. Paula [Woo82]. Pauli [All56l]. Pauling [Dow79a, Sha59, All60g, Ano56l, Ano63w, Don64, Kal60, Kal61, Pau70, TG70]. Paving [Lat79]. Pawn [Fin01]. Pax [Fis90, Sha84c, Szi60a, RK89a]. Pay [Hol95, Ano92t, Ano96o, Kap92, Kat90b, Pai83b, Rot92i]. Pay-more-later [Hol95]. Payback [Woo00d]. Paying [Joh65b, Med91, Cle87b]. Paymaster [Dew52]. Paymaster-General [Dew52]. Payments [Ols74a]. Payne [Hob85]. pays [Abb84b, Ano93s, Gar83b, Pen84, SG83]. PBS [AL79a, Bar90b]. PCs [Wol88]. Peace [Amd65, Ams76, And78b, Ang78, Ano52w, Ano55t, Ano60m, Ano74i, Ano76c, Ano79o, Ano88e, Ano89b, Ano89g, Ano92-32, Ano93v, Ano95i, Aus47, Bal79a, Bar77c, Bar78b, Bar79a, Bar80b, Bar81a, Bec62, Ben61, Bha03, Bla84, Blo66, Bor71, Bou66, Bre93, Bun79, Cha66, Che82, Cof62e, CS63, Com57a, Coo76b, Cor56, Dem60, DS49, Dou68a, Dow84b, Dul57, Dur69, Exe57, Elm68, F.61, Fed58, Fel68b, Fel68d, Fel77o, Fel82g, Fer63, Fin65, Fly02b, Fow55, Foz61f, Gar60, Got62, Gre82a, Gre07, Hig50, Hol96a, Hue77, HC83, Hun85c, Hut46a, Hux47, Jac62a, Kal77b, Kil62, Lal65c, Lap56d, Lau75, Lau64, Lei74, Ler46, Lor00b, Mah59b, Mar58b, McG62, McM50, McM58, Mel75, Min55, Mor85a, Mot49, Mow83, Mun65, Mur84, Nag81, Nat63, New67, NB77, NB79, Olu71, Orr78]. Peace [Pug77, Pan07a, Pau67, Pau79, Pea58, Pie58, Pom67, Rab59l, Rab69e, RP66, Rih65, Rot61, S.55a, Sha59, She56, Shi50, Sin63c, Spe55, Spi74, Sta79, Sti63, Sti48, Sto58, Sto59b, Swe86, Swe88, SM89b, SG59, Szi55a, Szi55b, Tel46a, Tel60, Tho80g, Toy84a, Ure46b, Ure48c, Wei68b, Wei03, Wu60, You82, You85b, Zar64, Zei64, vS78, Ack88c, Ack90f, Ano82c, Ano84l, Ano85p, Ano85i, Ano93w, Ano94l, Bar83a, Bar80d, BR78, Bou80, Bou82, Bro85g, Bro85h, Bro86f, Bro86g, Bro86h, Cor95, Cox86, Dav77a, Day77i, DW85, Don90, Eme76, Eva84, Eve89, Fel76e, Fel78e, For92, For20, Gan76, Goo85a, Gor96b, Gri91, Hal84, Han98b, Isa90m, Kal83c, Kat90a, Kel84, Loc92b, Mat83b, Moo95o, Nag84, Pug76, Rot94a, Sch92b, Sim84a, Sim84e, Sin79b, Ste82a, Suz84, Tay81, The84]. peace [TT83, Toy84b, Tru84, TM94c, Wei87, Yor88, ZGD83, Ano93t, MM82, Mar82c, Alc63, Ano62-28, Ano86i, Ano95x, FCC^{+}46, LC78, Nad98d, Sch69b, She61, Sto56b, Sto59b, RK87a, Bus85, Cha62, Div70, Kap61, Mye88, Nef78, Sha84b, Sto56b, Wal87a, Alc63, Ano81f, Ano86i, Cor94, Fin65, McL49]. Peace-building [Hol96a, RK87a]. Peace-Keeping [Fin65, Cox86, Div70, Fin65]. Peaceful [Ano54u, Ano54v, Ano56j, Ano60r, Bha55, Coc59, DR71, Don90, Fed70, Foz62b, Gla75d, Ing70, Lon76a, Pon60, Pre62, Rab68c, Rei50, Rei59, Sim54e, Ste60a, AJSB79, Hui91, Roc18, TBBS85, Myr75, Bie68]. Peacekeeper [Ack90k, Mti96b]. Peacekeeping [Bar69, Har99, Isa93o, KM90, Lal64h, Lal66h, Mar65d, Nea04, Hul98, Moo95q]. Peacemaker [Coh96]. Peacemakers [Kre89b]. Peacemaking [Ano84d]. Peacetime [Aeb48, Cha58a, Gom78, LS58, Ter58]. Peak [Ano08h, Gle24, Kam20, SCHD20]. peak-and [SCHD20]. Pearce [Pat81a].

Pearl [Mar58a, Ols62a, Smi59a, Ure45]. Pearson [Gol95d]. Peas [Isa00b]. Peasant [Ahm74, Bla97]. Pebbles [Isa01d, Ano90-34, Cha89a]. Peck [Ull63]. peculiar [Gre82a, Lan82a]. Pedro [Con88]. peek [KM89]. peer [Fin77a, Fin77b]. Pegram [Wol58a]. Peierls [Sha50, Fel86a, Lan57]. Peking [Que70b]. Pell [Pel64]. Pelted [Pri02]. Pelz [dF79e]. pen [Loc92b]. pen-chant [Loc92b]. penalties [Gol97]. Pending [Ano65h]. Pendred [PP63]. Pendulum [Tho80c]. penetration [Din78]. Pennant [Lew67b]. Pennies [Har99]. Pennsylvania [Ada73]. penny [Zag94d]. Pentagon [Gol77, Kap67, Mil86b, Rot96b, Sch71a, Ada86c, Ano91-49, Ano92-28, Ano94j, Ark87g, Ark95d, Bro87c, Fer87, Gaf86, Isa89g, Isa90x, Isa93f, IK01, Jac74b, Kis89, Kri86, Lal64b, Lau75, Lyn86, Mac92, O’S86, Pra05e, RL76, Rob91a, Sch74c, She05b, Wal89c]. Penycate [Dow81f]. People [Ben71b, Bro47c, Com50, Cut58, Das68, Dow82e, Due21, Ebe47, Foz61g, Gra89, Gus04, Hub57, Jen74, Kal77b, Kra91, Lon58, May47b, MK51, Mus04, Pau57, Rab54b, Rea80, Rie59, Sim59e, Wal07, Wil75b, Ano94o, Cha88c, Ebe20, Kam79, Lop90, Moo96f, Wat82, Ano80j, Ano91-40, Ash81, Esp72a, Bul07c, Zho79, Bre73, Fin57b, Ros90a, Wil80a]. People-Free [Das68]. Peoples [Ano54-30]. Pepeshinskaia [Dob54]. Pepsi [Ano91-35]. Percent [Ano92o, Ano93x, Eva84]. Perception [Das96, Gar74, Liv63, Wil81a, Slo12]. Perceptions [Bab93, Cra83b, Ekl80, Ler81, Orl83]. Percy [Hol62]. Peres [Coh96]. Perestroika [Dar91, Lar89, NA89s, Ack88k, Eri88a, Eri88b, Kap91a, Wei88, Cam89]. Perfect [Etz07, Kae53, Com76b, Fly98b, YSW21]. perfection [Ano94m]. Performance [Aur71g, Joh83b, Kom80]. perhaps [War14, JH16]. Peril [Bar79b, Hig66, Lap54d, Bab93, DeV95, Doy17, Mec20f, Mor83c, Smi65, Smi70b, Tsi83c, Wei79, Mec23d]. perilous [Mec18a, Pat88, Zag95d]. Perils [All50, DFJ^{+}63, GC83, Hul98, Ing70, Thi58, Cal17, Cav82, SAB^{+}16]. Period [Ano78g, Dje72b]. Perish [Kur79, McV71, Dro22e, STRS80, TR80]. permafrost [Dro25c]. Permanent [Nat63, KB80, Sto59b]. permanently [Man77]. Permits [Ano46r]. Permitted [Gol70a]. Peronists [Kes89]. perpetual [Yor75b]. Perpetuating [Sti78]. Perry [Hun87, Pri70, Mec19c]. Perseus [Rot99b]. perseveres [Coo85]. Pershing [Ano80i, Ark83d, Fel80k, Pai80a, Sha84e]. Persian [Hal93a, Ste86c]. Personal [Ame80, DuB48, Ehr72, Fal70, Mey08, Phe56, Yor70, Dan54a, Kap79, Sim81]. personalization [Gro15]. Personally [Gus07]. Personnel [Ano47-27, Ano49a, Bey46, GM54, Gre64b]. Perspective [Abb90, Ack86, Adi69, Ahn10, AGZ91, ABKS10, All68, Ang78, Ano83p, Ark84e, Ark92, Arn98c, AuC84, Ber80, Bal79a, Bal84c, Bal84b, Bea89, BT94, BP90, Bou80, Bun79, Com81, Cal97, Car83a, Car97, Coa88, CM91, Coh96, Col93a, Cra84g, Cro86, Day92, Day95, Dor90, Dun85, Dys81, Ear86, Eps87b, Eps88b, Fah90, FJ91, Fis91, Fra80, Fra87, Fri97, GSB^{+}83, Gar93, Gar97c, Gol98, Gol90b, Gom83, Gor84, Gra86, GK94, Gro81, Ham90, Har84b, Hay93, Her88a, Hia87, HRS81, Hol75, Hol84a, Hor88c, Hou72, IS92, Isa92m, Isa93n, Isa93m, Isa94e, Isa94d, Ist98, JPR90, Kap92, Kat80b, Khr96a, Kis93b, Kis89, Koe79, Kre85b, Kre91b, Kre92a, Kre95, Los57, Lea89, Lig84, Lin84c, Luc90b, Mal79, Mar87a, Mar87b, May85, MvH97, Mül90b, Nan78, Nat86a, Nat84, Nel92a, NB79]. Perspective [Obu89, Orr78, Osg67, Pop81, Pau79, Per92a, Plo86, Pop93, Pot97, Pow89a, Pra98, Rho90, RW80, Rot95l, Sal88, Sch92b, Sco84b, Sea89b, Sea61, Sha81c, Sha84e, Sha93c, Sha94, Sha95c, Sha97, She87, Sho92c, Sig87c, Sim91, Smi87a, Smi90b, Smi94b, Smi95b, Sør81, Spe89c, Sta79, Sta81, Swe87b, The79, The84, TS92, Tho83, Tho80i, Tim87, TM94a, TM96a, TM96b, Tru83, Tsi82c, Tsi83f, Tsi87, Tsi89c, Tsi90, Tur98, Vol74, WH89, Wei91b, WW80, Wei80c, Wei82b, Wio81, Woh68, Wol91a, Yaz88, Zac82, Zam93, Zuc82, vH75b, vH79, vH85, vH89, vS78, Cha11, Coo11, Fel79e, Flo78, Gre81, Iwa82, Jor11, LJ17, Mar78b, Mel78, Oel20, Sau10, Sha83, She94, Tre21, Wei78f, Fra83, Fre76a, Sco79]. Perspectives [Ber70a, Coh74, Fel84a, Hir88, Lal65f, Mul64a, Rab65e, Tod81, WS21, CK20, Ste85c, Ano78l, Ano86i, Ano62l]. Persuading [Hug87]. Persuasion [Hal93a]. Perturbing [McN76]. Peru [Ano70e, Rad92a, Rad92b]. Pervaiz [Das96]. Pervasive [Saw81a]. perversion [MK79]. Pest [vdB65]. Pestel [Han75]. Pesticide [dF79c]. pestilence [Kel98a, McL49]. Pests [Cha77, Wil73b]. Peter [Ano05b, Bar80m, Der88, Dow80c, Gol92a, Ham93, Her82b, Key81, Kre89c, Lev84b, LP92, Met83, Mil85c, Sag86, She79, Ano90-61, Dro20a, Bul08n, The90]. Peterson [Ano53v, Ano85d, Lap54a]. Petition [Kal82c]. Petroleum [Ano78m]. Petropavlovsk [Han91]. Pets [DIJ73, SB73]. Pfaltzgraff [Swe86]. Pfeffer [Gol69b]. Phantom [Lon95, Ros89a, Rot05b, Spa95b]. Phantoms [LKWR89]. Pharmaceutical [Ben69, Gil69b, Gil69a, O’B69a]. Pharr [Wil69b]. Phase [Ano82o, Jac74b, Nat55, Shi48c, Gla12, GP18, Lov13, Mat12, Sch12]. phase-out [Gla12, Lov13, Mat12, Sch12]. phases [Ros88]. Phenomenon [Mar67a, Mic69, Hen78b, Hil12, Rot92t]. Phil [Dow80h]. Philanthropy [Whe84b]. Philip [Ano62l, Arm68, Fay75, Fre84a, Gen84a, Gri83a, Hal93a, Kat52a, Kus83, Rab60b, Rab98, Spi72a, Sti83, Wei89b, WKW57, Wil77b]. Philipp [For80]. Philippines [Cha85a, Kes85a, Mor71, Mur70]. Phillip [Ano78f]. Philosopher [Ano15g]. Philosophers [Ell56]. Philosophical [Fra57, Kat54b]. Philosophy [Adl57, Fra57, Ley51, Pet63, Pol57b, Sch59b, Soy89]. phobia [Ark87g, Ark94c]. phone [Ano95i, Dro22f, Gar78a]. Phony [Rot02d]. Photo [Coo77, Lan90l]. Photo-Finish [Coo77]. photochemical [Bro76b]. photograph [Ano90j]. Photographs [Moo95d, A.52, FS95]. photography [Gal13]. photosynthetic [PW76a]. Photovoltaic [Kel79, Tri77, Wol76]. Phyllis [Rea80]. Physical [Ano45m, Bet75b, Bok48, Cog82, Kel79, Kow49, Lib65, Pal84, Rom90, Nat77]. Physician [Ben69, Bro66]. physicians [Ano85m, Fin77a, Str84, Gar83a]. Physicist [Ano22c, Art68, Bac48, Bor65a, Bor65b, Hol53, Kim87, Mas48b, McD74a, Seg70a, Ten75, WW59, Wei55c, Wil77b, Dro23b, Seg70b, Wil70b, vL47, vL48a, Wil71, All59l, LL90]. Physicists [Ano57h, Ano67n, BBF^{+}57, Rin52, Rot96i, Sei50, ASB^{+}84, Ano93u, Bir89, Cas84, CFG^{+}81, Hol84c, Rot95d, Mar79a]. Physicotechnical [Hol86a]. Physics [All50, All60h, Ano48o, Ano60f, Bon72, Bor58a, Bor60, Fal70, Fri59, Gou78b, Hei53, Hei70, Hei71, Jac71b, Kat54c, Lan56b, Loo50, Mar70, Mer46b, Opp48b, Ran50, Rid50a, Rub73, Sal65, Sal68, Wil77b, Hol94b, Jos87, Lan92a, SW59, Thi46, Wei78b, Ano50-36, HKL^{+}49, Wei56b, All56l, Ano56b, Fla68, Gol96b, Hei54, Her90b, Smi61, Ano91r]. Physiker [vL47, vL48a]. Physiology [Gan52, Wil73a]. Piasecki [Abb84a]. Picadon [Dow83f]. Pickering [Bul08s]. Pictorial [Ano46t]. Picture [Ano78k, Ano94q]. Pictures [Joh81a]. Picturing [FS95, Moo95d]. Pie [Ark84d, Aue03a, Coo90]. piece [AH92b]. Piecemeal [Fel75b]. pieces [AZ89b, Bey12a, NA91i]. Piecing [CRB^{+}09]. Pied [Elm76]. Piel [Sim62]. Pierce [Dow83e, Lap57, dK60]. Pierre [Gal54, Rot62, Sei62, Bar73a, Gan82, Nie55]. pig [Fin92]. Pigs [Ise01, Tho80g, Ack90e, Dow82d]. Pika [Mes83a]. Pilcher [Dow84d]. Pile [AT62, Ano46i, Ano46-58, Ano48-27, Ano49r, Wat74]. Piling [Abb04a]. pill [Ano90-31]. Piller [Che88]. Pillory [Lil48]. Pilot [Gol11]. PIMBY [Shk92]. Pimping [Lef76]. pinpoints [Rot93w]. pioneer [Ano93-29, Coh96, Eav18]. Pioneering [Rab67e]. Pioneers [Mor80, Wal74, Wil74a, Wil75d, AC02, NA92k, Qua80]. Piotrow [Rea80]. Pipedream [Sew67]. Pipeline [Jac74b, AZ90, Rot93o]. Pipers [Elm76]. Pipes [Bla85a]. pirates [Moo91a]. pistols [NA87e]. Pistons [SC75]. Pits [Pai03, Ano96m, Buk97c]. pittance [McN77]. Pitted [Nel74b]. Pity [Ano91-40]. Pitzer [Ano49-32, FDC52]. Place [Beu49, Coh02, Dow83e, Fly00b, G.84a, Ham83, Lib65, Rep03, Smy50a, AZ89b, Ano95m, Bli76, Hyd91]. places [Ano93r, Ano97j]. Plague [Kel98a, Sch05f, Tuc97, Wei66d]. Plagued [Mar06b]. Plagues [Paj52]. Plain [Ber73b, Fel79h]. Plame [Mec16b]. Plan [Ano46o, Ano46-47, Ano46-78, Ano49q, Ano55t, Ano60f, Bli70, Cav50, Fis62c, Gar00b, Gro48, Han46, Har61, Kar88a, Lal64i, Lev84b, Lon73, McG84, McM48b, Mue69, Myr68, Nea61, Rab47k, Shi50, Sim54d, Sme74, Tay74, Wil72b, Zei50a, Agn77a, Ant78, Cav79, DSL^{+}87, Gut93d, Gut93c, Hol95, Kau90, Lac90, LL12, Mag88b, MDB81, MvH97, Mit58, Nas18, NA89m, Noy47, Nye77, Pai82i, Sim87b, Smi96, Wur92, Pai81a, Mas48a, Pie19, Prz82]. Plane [Mat98e, Kel98b, Rot92j, Rot97c, Sie07a, NA87d]. Planer [Hel85]. Planet [Dow83b, Eng08, Key66, Lan58, Mea72, Moo03b, Pie55, Sim57a, Ano85i, Ano13a, Eav17a, Kre88a, Owe79, Tha63, Ano86i, Ven80]. Planetary [Ano67k, Ber81]. Planets [Ric71, Tur73, dV61, Ano78n]. Plank [Rab56m]. Planned [Ano49-29, Ber49, DIJ73, Mac67, Pol49, SB73, Fer53]. Planners [Ano53n]. Planning [Ano53n, Ano54x, Ano59c, Cav52, Cla77, CR71, Fox48, Gué69, Hil60, Hop49, Hur77b, Kel64, Kna63, Lov74c, Mon53b, Nac69, Pat81a, Ran50, Ric03b, Shi47d, Sym50, Bey80b, Bor14, Bru82, Cra82d, Gei17, HBD17, Hen78a, Lam83, Mec20g, Mor77, Sch54b, Rab52f, Kon67]. Plans [Ano46h, Ano46-72, Ano46-73, Bra49, Hek60, HK09, Por51, RR07, Wei55a, Alb84, Bur87, Cra98, Dea88, Gol90c, Zha15b]. Plant [Baw60, Cha68b, Gol11, Pie22, Sch74b, AH91h, Ano87j, Gar92, Gil18, Kom80, Lan90o, RS88a, Wil77a, Jas81, Lan58]. Planting [CPV01, Fie19, Fre63]. Plants [Ano49-29, Far71, HK09, Lym08, ML75, Rog71, Ano93-35, Cap77, PR85, Ram86b, Wil86b, vH79, Met83, Ing79a]. Plasma [Jac71b]. plastic [Ano94o]. Plastics [Smi94b]. plate [War14]. plateau [Mor82b, JM83]. Platforms [Ano48h, Rab59k]. Platt [Div79, dF79h]. Plaudit [Ewi60]. plausible [LJ17]. Play [Bar79d, Dow80f, Isa96f, Mar69, O’H01, Pra05e, Rab51h, Ust96, Ano82m, Ano96j, Ano98l, Ano15c, Kap91b, Kre88d, Lor92, Mia98, Pil82, Rob56, RH17, Sap80b]. play-by-play [Kre88d]. Played [Abe74, Gla75a]. Player [AH03c, Fri78]. Players [EO00, Gol88a, Spe89b, SB91]. Playgirl [LSB74]. Playing [Ack88e, Ark85e, Cav06, CO14, Fly00d, Gin76, Isa92j, Isa97b, Joh89, Ric75a, Rot98i, Sto76b, Wei53c, Dor85a, KM18]. plays [Wio83]. Plea [Ano51w, Ein48b, Ano94p, Lop76, Mak94a]. Pleads [Git74]. Please [Ano94n, Hit01, Pra03c, Ano95u, Cai90, Fer10, Mak95b, Moo98c, NA90d, Osa77, Sør77]. Pleasure [Pre75, Pre76]. Plebiscite [She85c]. Pledge [ABB^{+}50, Fel67d, Len73, Ano86d, KW86, Moo97q]. Pledging [Dug91]. Plenipotentiary [Rot93j]. Plenty [Cav04, Eva63]. Plight [Mey56, Rab72a, BP85]. PLO [Coh93b]. Plot [Gre64a, Boy89a, Kat90a]. Ploughshares [Mec21d]. Plowing [Ano14f]. Plowshare [Pre62, Ano95m, IS67, Joh70, Ste60a]. Plowshares [Pol91b, Day77n, Mag16]. Ploy [BK03, Lon87, Pes87]. plugged [Sch75e]. Plugging [Rot93o, Rub74]. Plunging [Isa03a]. Pluralism [Gre73a]. Plus [Bal54]. Plutoburg [Isa92i]. Plutonium [AZL90a, Ano49n, Ano94u, Ano97n, BM02, BDF^{+}92, BP90, Bol88a, Day76c, Fel75c, Fin92, Gee71, GC77, Hib94b, Hib94c, Hib94d, Hib01, Jan15, Lan89f, MB99, MvHKN01, MC75, Mil85a, Moo97m, NA99e, NE76, Pat82b, Pen77a, RR15, Ric77, Rot93a, Rot01d, Rot01g, STC74, Bul02f, Wal21a, WPB84, Zha15a, Zim81b, AH92h, Alb94d, Ano84p, Ano93-35, Ano03i, BP84, Dah77, Day77d, DeV95, Eds76, FTvHW76a, FTvHW76b, Hib96a, IPB84, Kün96, LC94, Lym96, PB84, PWS^{+}96, Pat88, RS88a, Rot77, Sal80b, Sto90, TN90, TvH17, Tra24, Zha20, Ano13c, Lym97]. PNEs [LT99a, You77a]. Pochin [Lin85]. Pocket [Ano78n]. Pod [Isa00b]. Podvig [Nor02a]. Poet [Ols59]. Poetic [McI69]. Poetry [Sim65]. Poets [Ano66j, McC65]. Poewr [Fra83, Whe84b]. Pohl [McC88a]. Point [Abb04c, Bet69b, Fry56d, Gom75b, Han75, Ing55b, Pan54, Stu48, Los54, Ump75, Ada86a, Cra86a, Dro25b, Ima93b, KGS25, NA92e, Spe18, Ano89r, Bey80a, Bin54, Dan54b, Fra54, Sti83, Bea86a]. Pointing [Fie78]. Points [Com75g, Rab50d, Rap65a, Wei06, Ano91-34, Ark98f, Dro25a, McK25b, NLS25, Wei82b, vH76b]. Poised [Dro21c]. Poison [Lyn78, Mes71, Rot04a, Spi88, Tan88, Gao17, Smi92b]. Poisoned [Dan90b]. Poisons [Rid50b]. Poking [Gri91]. Poland [Pol68, Rab67h, Kih47f, Prz82]. Polanyi [Ano87-32, Roc81, BB52, Pol62a]. Polar [Fle70]. polarization [Pat79e]. Pole [SJ68]. Police [Fed58, Hib94b]. Policeman [Mas67, Isa93p].

Policies [Ano52c, Ano87n, Ano89e, Cof65b, Dad67, Dre82b, Fel83d, Fry65, Gra66, Ham61c, Kat50b, Lar65, Lov74b, Mor85c, Pol61, Afh88, Cle11, Dor96, Fel80l, Tos82, Toy85, Wan83, Wol80, Cra84b]. Policing [Rot97b, Ure61]. Policy [Adv62, AS62, Ahe91, Ahm74, Ali57, Alp62, And65, Ano45n, Ano46x, Ano49p, Ano51p, Ano53-46, Ano57c, Ano62v, Ano63z, Ano66m, Ano67b, Ano69j, Ano84h, Ano86l, Ano87q, Ano89f, Ant74, Aro51, Aro52, Aro55, Aur71e, Aur71f, Aur72e, Bea67, BH62, Bos51, Bro48, Cou57, Cha69b, Dad71, De 74, Dea53b, Ded59, Dou68b, DuB46b, Eme48, Edg52, EBU^{+}52, Fed55, Fal68, FGJ65, Fei60, Fei61, Fer03, Fis61a, Fit66, Foz62a, Fre71, Ful69, Gal56, Gil68, Gol67b, Gor60, Gor51a, Ham61b, Han98a, Has67, Het55, Hil65b, Hit01, Hou72, Hum58, HM72, Hun87, Ing52a, Isa04a, Jac79, Kah50, Kal87, Kar72, Kas69, Kes85a, Kil59, Kin54, Kor72, Kov68, Kri85, Lal64b, Lal66i, Lal67h, Lan70, Lea69, Lee74, Lew67a, Lew67c, Lin69, Mac67, Mar75, Mar48d, Mas48a, McK74]. Policy [McM65, MF63, Mea69, Mer90, Mes70, Mes84b, Mes84a, Mor67, Mye87a, Nat55, Nea63, Nit57, Noy57, Oom46a, Oom46b, Opp53, Opp63a, Pre48, Pan54, Pan07a, Pot53a, Pot53b, Rea68, Rie59, Roc83a, S.46, Sen57, Stu47, Sca91, Sch71b, Sha93a, She61, Shi47a, Sim54d, Sko68, Sla88, Som91, Spi75, Bul09e, Sta60b, Ste57a, Ste62, Sti60, Swe89a, Ter04, Tod79, Tuv53, Voh70, Wal76, Wei46, Whe51, Whi71, Wie64, Wil59, WG69, Wil72b, Wri63a, Wri63b, You64b, Zuc52, dF79h, dF79b, Ald17, Ano86t, Ano14e, Ark95d, Bal84a, Bro85a, Chu76, Cla81, Cle92, CM12, Dal83, DM85, DM24a, Din78, Dje86, Doy53, Dro19, Fel81b, FD78, Fri88, Gap86, Gap90, Gol78b, Gra81b, Gre12, Hal89, Hec16, HC87, Hir76, Hof86, Hol77, Hou82, Isa92l, Isa93e, IS15, JOO15]. policy [Lak78, Lan82b, Lap76, Lym11, Lyn88, Mac77b, Mcc88c, Mez12, Mot82, NH08b, Pap98, PY98, Por82, Ram88a, Ram13a, Rep89, Ric76, Rom91, Sco08b, See79, Sha85c, Sie82, Sok87, Bul09h, Tak11, Tho12, Vog87, War14, Zik21, Ano85d, Cro65b, Fra83, Her83a, How65b, McK74, Ano57d, Ano60f, Bla87a, Cra84c, Dow88, Fel73a, Fis83, Gol69b, H.84a, Har79b, Hob85, How65b, Mes64b, Moo95b, Nel74a, Pal64b, Pat81a, Rab60b, Sch69a, She83b, Sim85, Sim54a, Tod79, Wig62b, Woo82, You66c, Zim81b, Fre88, Tod78b]. Policy-Making [Whi71, Mes64b]. Policymakers [Gol88a, Ede89]. Poliovirus [SW10]. Polish [Ano48-33, Ano91-51, Cav90, Lal64i, Zie77]. polite [Ken85]. Politeo [Dow82g]. Political [AS00, Ano49-30, Ano69q, Atl54, Aue78, Bro79b, Bro67b, CH71, Cra83b, Dan70, Day75c, Der88, Der69, EO00, Fel83a, Fil76, Fis83, Fly99b, Fre84a, Gen84b, Gle94, Gol95d, Gre67d, Hig66, Hol86b, Hol03, Hut53a, Isa00e, Jan64, Keu83, Lal65a, Ley51, Maz69, Min55, Mor54, Nad98b, Nye82, Ost67, Pal64b, Pri74, Rab69g, Rab72f, Roc81, San66b, Shi46b, SK69, Smi75c, Soh64, Sol70, Ste85b, Str81, You68c, ZS93, Zim48, dJ57a, Alb83, BV98, Chi17, Cor78, Gol78b, IK21, Kla76, MD24, Nea78a, Nea78b, Osg83, Pai82h, Wol17, Spe95b]. Political-Economic [Bro67b]. Politically [Ark99e, Mec21f]. Politician [Pya74]. Politicians [DSL^{+}87, Hal56, Shi54d, Wal87b, Cah76]. politicizing [Bou98a]. Politics [Adv60b, Abb91, Ame66, Ano60o, Ano60n, Ano60p, Ano60q, Ano66m, Ano69n, Ano85d, Art68, Bar73a, Bar79b, Bin54, Blo63, Bol51, Bon72, Bor60, Bou66, Che70, Cle74, Cof62e, Col84b, Con81a, Con88, Day75o, Dew52, Div70, Dow80h, Dun48, Eds62, Eis68, Esp72a, Eve85, Fal70, Fay75, Fra82, Fre60, Fri72, Gan82, Gen86, Gla62c, Gol93d, Gol49, Gou78b, Gre73a, Har64a, Jas81, Kas65, Kel98a, Lar62b, LLH^{+}62, Lee74, LT70, LP92, Mac68a, Mac66, Mar78a, Mar64h, Mar61a, Mer46b, Met83, Mil47, Mor66a, Mor66b, Mor85a, Mye89, Nan73, Ney69, Oli58, Pag68, Pal64b, Pat81a, PR62, Pie73, PvH72, Rab56k, Rab63i, Rab72c, Rav65, RMK65, Ric75a, Sar85, Sch68e, Sch68g, Sch69d, Sco79, Sil64a, Sim60b, Sin05a, Sno68b, Ste76a, SG64, Tel47d, TS01, Til64]. Politics [Til60, Tow66, Tra90, Tuc97, Tum87, UH56, Van91, Van88, WW50, WW59, Wol63, Woo82, Woo73b, Wor63, Wri63a, You67a, Yud83, dF79c, dF80, dJ63, Aja78, Ano80j, Ano90y, Bar78a, Ber82, Dje86, Dro16b, EE81b, Fel77c, Fel78g, Fel79n, Gal81b, Gre82a, Ham14, Hib94d, Jas91, Kan11, Ken81, Ken87c, Kor77, Lan85, Lan82a, Mec21d, Mow83, NK20, Pat86a, Ros78b, Sch12, Sut96a, Tak11, Zho79, How65b, Mor62d, Wri63a, dF79d, Cah94, Cor94, Fre76b, Her85, Mor62d, Pal02, Rea80, Sch68f, SB91, Swe86, You65a, Bro79b, How65b, Kin80b]. Politisches [Leb91b]. polity [Zim78]. Polk [Cro65a]. Poll [Ano99z, CD48, Sta48, Ano86p]. Pollak [Wal83b]. Pollard [Wel79]. Pollock [Roc85]. Pollutants [Lee73]. Polluted [Ber70a, Dix65]. Polluting [Fre73]. Pollution [Ano70g, Ant69, AG91, Dow80d, Fer69b, Gol72b, Got69, Gri65, Gus70, Hil65a, Jac73g, Jac74d, Kne65, Lee73, LR71, Lov68, Mid65, New73, Ney73, Ram70, SEM70, Sch70a, Wol65, YY74, AM21, Fie19, Ham91b, HNB81, Pat75, Jas81]. Polo [Ols62a]. Polsky [Ano09]. Polyakov [Gar00a]. Pomerance [Ano80l]. Pondering [Wig61]. Poo [WM78]. Pool [Ano78p, Lav55, Rab54e, Str54a, Tat78]. pooling [For92, For20]. Poor [Nas62, Rab68f, Sal63a, Gol91a, GB77, Kap76, Nye79, Sch80d, Nan73]. Pope [Pop52, Pep15]. Popkens [Smi83]. Popkin [Ale69]. Popping [Gou23]. Poppy [Nas04a]. popular [Wal90c, YSW21]. Popularity [Isa03a]. Populated [McE88]. Population [Ahm74, Ano66o, Ano86-27, Aur72e, Bil86, Bra74, CF74, Cho86a, Cho58, Cho86b, Coo56, Dal86, Dar58, Dav86a, Dje86, Eck78, Ehr82b, EE86, Eng08, Gid73, Har72b, Hau64, Hoa64, Mei52, Mol62b, MM86, Myr72, Nor74, Not51, Orl66, Orn67, Rab51m, Sax51, Sch54b, Tae51, Tod79, Wie83, You71, dN71, Str86, Woo81, Eze16, Gra16, Mye87a, Wan16, Ano62l]. Populations [Bio73, Car55]. Populi [Gom62b, Gom62c]. porch [Sti18]. porcupine [San76b]. pork [Har86d]. porpoises [Rot97j]. Porro [Div79]. Porter [Bus85]. Portrait [Dow82c, LP97, Sch80c, Siv75, Ber04c, MM11, Leb91b]. Portraits [For80, Gus04]. pose [RK19]. posed [Tra24]. Position [Ano50q, Ano60r, Arm53, Bar10b, Kib10, Kib12, Kor10, Lan10, LW61, Nay12, Pau67, Pic10, Ram10, Shi49, Squ12, Szi47a, TB70, Tra72c, Ano91-49]. positioning [Ano91-49]. Positive [Hug56, Hux47, Pat79f, PBG55, NLS25, Roc97a]. Possession [Fry55]. possibilities [DL21]. Possibility [Lar65, Lwo60, Sew67, Sha69, Szi60a, Tor19]. Possible [Ano60a, Chi59b, GLMS60, Sel59, Arb19, Bad78, Buk97b, BR15, Gar84a, Jer85, Led94, PR14, Tak17]. possibly [LvH18]. Post [Ano55f, Ano78m, Ano97m, Bre93, Bro67a, Bru82, Fel14, Gro54a, Gro54b, Gro54c, Gro54d, Ise92, Lea69, Cra82d, Dro16a, Hol96a, Lym11, Suz11, Tak11]. Post-Apollo [Bro67a, Lea69]. Post-Cold [Ano97m, Hol96a]. Post-Fukushima [Fel14, Lym11, Suz11, Tak11]. Post-game [Ise92]. post-Nobel [Dro16a]. Post-Nuclear [Bre93]. Post-Petroleum [Ano78m]. Post-war [Bru82, Cra82d]. Poste [Ano13b]. Postel [Roc85]. Postmortem [Gar84c]. Postol [Gre82b]. Postponed [Ano46-64, Rot77]. Postponement [Mur51]. Postscript [Aki05, Gal78b]. Posture [Ark94a, Dvo07, Isa02e, Rob20, Bla20]. Posturing [Ark01a, Kri02c, Kri02b]. Postwar [Ber86b, MB86, She83b, Szi47a, YG77a, Ski53, Cav61]. potato [Ano95i, KKL15]. potatoes [Ano94n, BP95]. Potential [Alb89a, Ber72a, Hir73, Mac97b, Tay87, Ban11, Coa47, Ing76, Pal77a, McK74]. Potentialities [Dub60, Coh76c]. Pots [Oli66]. Potsdam [Hig66]. Potter [Ano86k, Fra83]. pound [Ano95n]. Poundstone [Gai92]. Poverty [Dev69, Dix66, Fel66d, Kam71, Key66, Smi74c, MS77c]. Powder [Dje75, Hib94b]. Powell [Rab70a]. Power [A.55, Abb85, Abr73, Ack09, Ahm99, Ano45a, Ano46i, Ano46-39, Ano46-71, Ano46-74, Ano48p, Ano48-31, Ano49x, Ano51l, Ano52t, Ano53d, Ano53k, Ano54t, Ano54u, Ano57b, Ano85d, Arn80, Aro60, Bas80, Ben71a, Ben71b, Ber53a, Ber56, BAA^{+}75, Bet77a, Bet77b, Bro16, Brü55c, Buc60, Cav47, Cis52, Coh53a, Com74, Com75a, Com53, Cra83b, Cra84b, Cre70, Dan46, Dea53b, Deu60, Dow81g, Eli71, Eme63, Far71, Fel55, Fel74b, FL96, FK71, For09, Foz63b, Fra83, Fri54, Gal91, Gan83c, Gar72, GA74, Gil83, Gil63, Gof71, Gol90d, GM09b, Gou47b, GR62, Gus70, H.84a, Haf49, Hol77, Hot76, Hus04, Ing46, Ing75, Ing79a, IW49, IW52, Jac73b, Jas81, Kat92, Kat53a, Kno60, Kor72, Lam80, Lan08, Lan56b, Lap56c, MCAH07, Mal63, ML75, Mar46a, Mar46b, Mar54b, Mar98c]. Power [Mas76, McA76, Mea93, Men48, Mer79, Mes84b, Met83, Mus04, Nan73, Nel84, NA88m, Ola82, Par46a, Pau05, Pit52b, Pol69, PW76b, Pos71, Pra05e, Pry48, Rab62a, Rab72d, Rad53, Rea76, Rog71, Sac64, Sac84, San66b, Sch74b, Sch47, Sch49, Sch65b, SA62, Sha90, Sim53, Sim54d, Sne57, SC07, STC74, Spi72a, Spo53, SB91, Sto68b, Str07c, Sui71, Thi52, Tho80g, Til64, Tum87, Ust96, Wal07, Wal99, Wea79, Web06, Wei53b, Wei06, Wes99, Wes04, Wil67b, Wil72b, Wio83, Wit04, You66c, Zen76, Zuc52, Abb12, Abb89, Ack89e, Ack90j, Aga79, Ano79j, Ano80q, Ano83c, Ano87j, Ano95n, Ano14g, Ano15c, Ano15e, Ant76, AG91, Bei82, Bro76b, BHK^{+}21, Cam90, Cap77, Car94b, Coc81b, Coh76d, CEE85, Cor94, Cra82c, Dav77a, DJ13, DK80, Dro21b, Eme76, Fai80]. power [FTvHW76b, Fla87, Flo76b, Flo78, Fri97, Gro78c, Gil18, Goe79, Gol90e, Gre82a, Gri87b, Her84, Hol78, Hol80a, Hol83b, Hor89, HK13, Hum83, Ibs87, Jas91, Jas24, Jor18, Kas15, KA90, Kop15, Ler81, Lev12, Loc14, Loc94, Luc90b, Mar98b, McN82, Mec16f, Mec17b, Mec23c, Mir89, Nat81, NP80, Nel81, NA87a, Pat90a, Pie22, PR85, PW76a, Pos24, Pow83, Pre80, Roy76, Ram11, Ram13b, Ray80, Reh14, Ros85, RS16, RH17, RF21, Rot94d, Rot97i, Sak84, Sch13b, Sha96, SR81a, SL80, Smi95a, Spi83, Squ21b, Ste80b, SHT^{+}87, Swe77, Wal90d, Wal21a, Wei86a, Wel80, Wel16, Wes98, WF90, Wil77c, Wil80c, Wil80d, Woo12, Woo97d, Zel76a, vHW76a, Pit52a, Lan82a, Abb91, Gol67a, Jer87, Lof51, Lon78, Per94, Fla80b, Her83a, Sto55c]. power/nuclear [Goe79]. Powered [Ano55g, Mar00]. Powering [NA91b, NA93l, Rot01e]. Powerline [Mor82a]. Powers [Cof67, Dia83, Iri79, Lew70d, Rus52b, Swe93, Zlo47, Eps97, NA90f, NA90g, Lip61, Rei74]. pox [Lev98]. pp [Ano10a]. PR [Ano91-42, Arn91, Isa91j, Kre86e]. Practical [Gla62a, Kal53, Kil65, Kis66, Leb85, McG84, Mec16e, Mel53b, Rin57, Sew67, Wol55, RG13]. Practicalities [Dje72a]. Practice [Cah94, Eis72, Eis90, LS58, BS12, IS15, Rin84d, Ros87b, Tic94, Cra83b, Nei82a, Wri67]. Practices [Wil66, Luc82]. Practicing [Dje74]. Prados [Ack95, Cra83b, vL92]. Praetorian [Ark01b]. Pragmatic [Tsi78c]. Pragmatism [Isa99e, Zac76]. Praise [Ark99b, McI69]. Praying [Dro24g]. PRC [FD78]. Pre [Fra61b]. Pre-Atomic [Fra61b]. Preaching [Ano13e]. Precede [Boy10, Ram64]. precious [Moo92k]. precisely [Bar77b]. precision [KK98a]. precision-strike [KK98a]. preconceptions [Ano90-31]. Predicament [Par88b]. Predictable [Dys75, Mar88]. Predicting [RK90]. Prediction [Sac64, Woh64]. Predictions [Rot01f]. Predicts [Ano47n]. Predoctoral [Ano50b]. Preempting [Uhl04b]. Preemption [Kam05, Pai04d, Ram06]. Preemptive [Kri02b, Kri02c, PB92]. preface [Kap61]. Pregnancy [Bru57]. Prehistoric [Haw73]. Preliminary [Com48a]. Prelude [Bai75b, Dad71, Stu69b]. prematurely [Fel77g]. Premier [Khr61]. premises [RGK95]. Premonitions [Sea85a]. Prentiss [Ada52]. Preparation [Ano45c, Elm68, Zif53, FWCL20]. Preparations [Sub68]. Preparatory [Smi93c, Zer13]. Prepare [Wei68b, MD18, PFW18, SCHD20]. prepared [Ano92s, Hau17, Van20]. Preparedness [Ano91-50, Hoa57, Mat57, Ger21, Khr10, Sch16, Set16, Sim16]. Prepares [Ano46-61, Jac72a]. Preparing [Abr83, Ant50, Ark85f, Lei74, Mor80]. Preponderance [Per94]. Prerequisites [Sto59b]. Prerogatives [Fel83d]. Prescription [Wal81a, Bul09h]. Prescriptions [Ano85c]. Present [Ano46-75, Hir72, Kat75, Lap70, Lei72b, Mar48c, MD64, Nov77b, Pal55, Vol74, Wat51, Ano93s, Cra82a, Dou24, Fel78j, Fel80a, Ric98a, Rot77, Bre93]. Presented [Ano48t, Gen59, Thi46]. Presents [Ano46-31, Nic54b, Nic23]. preservative [Som83a]. Preserve [Col74, Rho07, Sag83]. preserved [Rho15]. Preserving [Hen17, Pod24, Rot93p]. Presidency [Wei89a]. President [Ano80k, Bou02, BWS^{+}80, Cap98, Gol78a, Her88b, Hor77, Isa92l, NH08b, Sok16, WH89, Ad 90, ADH^{+}73, Ano46-76, Ano46-77, Ano47o, Ano47p, Ano48-34, Ano50-37, Ano50-39, Ano50-38, Ano52s, Ano55q, Ano58d, Ano62w, Bal84c, Bar46b, Bla20, Bon53b, Buc79, Bun98, Eis54b, Eis54c, Eis58, Gen86, Ger21, Gol21, Hor77, Isa00c, Ken61b, Ken61c, Kor21, Lea47, LPT^{+}54, Lon73, Man21, Off02, Phy80, Rab54h, Saw53, Str56, Szi61a, Tay74, Tru46, Tru48a, Tru48b, Tru49, Tru53, Uni78, Wal46a, A.52, Hym91, Tho80b]. Presidential [Tru51, Woo92, Bel20b, Fie24, Kre88c, Hec16, Mem80, Mec16e]. Presidents [Ros90a, Kas75]. Presidium [BS57]. Press [Ack97, Ano45d, Ano58h, Ano90-43, Cam54, Fri56, Ken60a, Lew49, Mar55a, Nor88, Nor02a, Nor06, SB91, Str05f, Wei85b, Wei89a, Wil71, Ano91-33, Ano91-51, Arn91, CW81, Gut93e, Ano53-39, Hog68]. Presses [Eps88b]. Pressing [Isa88a, Smi21]. Pressure [MM77, Pal64b, Cor94, Fie21a, Roc97a, Wei82b, Dow80h]. Pressuring [Cas85a, Mac86a]. Prestige [MY67]. presume [Ano97j]. Presuppositions [Bur57, Man57]. Pretoria [Que70b]. Pretty [Rot04a]. Prevail [Rab62f, Las62]. Prevent [Rus58b, Top58a, Rus20, Tin23]. preventable [BB20]. prevented [Che52, Cla77]. Preventing [BB20, Fis81, Hia80, Kal64, ML09, Sch81c, Hen17, Wal81b, Ano91s, LaF90a]. Prevention [Ano87p, Bet65, Dre89a, KND10, LH05, Men64, Rus46, Wri47b, Ano96n, Cla77, FWCL20, Lif80, Pon82]. Preventive [Bal50, Voh71c]. Preview [Ano78m, Mat79]. Previews [Rab48g, Rab63d]. prey [Isa89k]. Price [AO95c, Alc63, Ale69, AC49, Cut75, Fel74a, Fer63, Gol61c, Kal62b, Lap59i, New73, Rea76, Sch80e, Siv80, Sto56b, SG59, Wal87a, Wu99, Zag94e, Ano92t, Bri80a, EF78, Fel79j, Isa89l, Lak82, Ras80, Sch92b, Win79, Fel84b, Gen84b, Opp62a, Bro81]. prices [Cav13]. pride [Ano96n, Chu94a]. priesthoods [Ano92q]. Priestly [Dow80j]. Primack [Day75c]. primacy [Tru85]. Primary [Ada47]. Primate [Rab73c]. Prime [Isa00d, RI83, Tru83]. Primed [Bla07]. Primer [Haf05, SS05, Sig10, Agn77b, AF88, Alb93d, BWL21, Hal89, Veg98a]. primers [Ano98h]. priming [Sea85b]. Primrose [Cob78]. Prince [Mos91, Wri63a, Dro16b]. Princeton [SB91, Ano48-35]. Principle [Her83a, Isa99e, Jon97, Kre87c, Moo96k, Sta00, Pon60, Kel79]. Principles [Ano47q, Ano72h, BS12, Fer53, Gan52, Mil85b, Mor85c, Sel69a, Wri67, Ano95a, Sta87]. Pringle [Her82b, Lev84b]. print [Ack89a]. printing [HV18]. Priorities [Hay78, Hek60, Mat75, Mun67b, Smi87a, Ano86-31, Ark98e, Isa88c, Kin84a, Lin84c, Sha81b, Gol88g, Gra84b]. Priority [Dje68, Fin53, Sak79, Wei80c]. Prisoner [Ano93w, Gai92]. Pritchard [WW62]. Privacy [Shi56b, Leo19]. privat [For80]. Private [Ano52t, Ano53a, Ano53b, Ano53c, Ano53d, Ano53-43, Klo85, Lil49b, Rie59, Ski53, Spo51, Bul08a, Tuc72, Wei53a, Dro22f, Hor77, Sam22, Whe84b]. Privatizing [Har04]. Privilege [Fel66d, Key66, LL90]. Prize [Ano63w, Shu81, Ano55k, Ano87-32, Ano92-32, Aur71c, CCD^{+}52, Don64, Dro16a, Dro23a, Fel82g]. Prize-winner [Dro23a]. Prize-Winners [CCD^{+}52]. Prizes [Kat52b]. Pro [Cam60, DRM^{+}09, Joh04, Rab59d]. Pro-American [Joh04]. Probabilities [Dje72a, Gre12]. Probability [Llu99, Ney73]. Probably [Sal16]. Probes [Kel61]. Probing [LH70]. Problem [Ale64, Ano46-106, Ano55r, Ano60f, Atl54, Bos06b, Car55, DeM76, Ded70, Fel74b, Fow55, Fre77a, Gol66, Gra61, Haf47, Har47, Hol74, Ing52b, Ing59b, Kar80, Kee76, Kra99, Lak64, Leg58, McC09, Myr69, Nac65, Nac69, Nas04a, Osb49b, PK63, Sas06, Smi58c, Bul09d, Stu69a, Szi55a, Ure49b, Ure61, Wes76, Wil75b, Wou71, You75, Zei50b, Alv14, DJE^{+}82, Due21, Gib91a, Klu22, La 88b, LBCA23, Lei89b, Mor87a, Rof19, Ros78a, Rot98f, Rot98h, San14, Smi77b, TvH17, vV23, Spe49, Spi49]. problematic [Hib96a]. Probleme [RK87a]. Problems [Ano45a, Ano48e, Ano49-37, Ano54h, Ano56g, Ano86k, Bat73, Bet70, Cha53, Che56, Com57b, Dar58, Dos72, Dre82b, Fav52, Fly06b, Gen59, IS67, Jan50, Kne65, Kuz58, Lan02, Lap57, Lap59a, Law85, Lew70e, Lon51, Mar65h, Ma’52, McM58, Mid65, Mil48a, Olu66, Orl66, Pla64b, Pol53, Rab54d, RK87a, Sto58, Tas51, Tra73, Ure69b, Wee73, Wie70, Wol65, Zei50a, Zuc52, Ano82q, Cim84, Con81b, Dah77, Hes78, Kap81a, Kre85b, Pow82, Rot93w, Rot95a, Tud21, Wei17, vH79, Ste76a, You66a]. Procedure [Ano48a]. Procedures [Cle48a, Cle48b, Fel83d, GL48, MB48, Mor53b, Sci48, Spi49, Wei52a]. Proceed [Bar81d, Ids81]. Proceedings [Ano47r, Ing52c, Lie46, Ano60c]. process [Her78b, McL11, Pan86b, Ano86s, Ull63]. Processes [Ano49n, Mid73]. Processing [rWeKHF47]. Procurement [Ano50-42, Kat50b, HSJH18, LST10, SL13, Tsi80]. Procurers [AH92j]. prodding [Moo91d]. Prods [Ano46j]. Produced [Mul55b, Ano81f, Ano90-64, Bar80m, Bar90b, Bar90c, Dow79b, Dow79e, Dow79d, Dow79a, Dow79c, Dow80e, Dow80g, Dow80f, Dow80c, Dow80a, Dow80d, Dow80h, Dow80j, Dow80i, Dow81g, Dow81f, Dow81i, Dow81b, Dow81e, Dow81c, Dow82d, Dow82c, Dow82f, Dow82b, Dow82e, Dow82g, Dow83b, Dow83c, Dow83e, Dow83a, Dow83f, Dow83d, Dow84d, Dow84b, Dow84a, Dow84c, HD82, Lad81, Say84, Sch80c, Wea82]. producing [AP88, AH91g, Ant76, Rot77, Wag88]. product [Lon76b]. Production [Ano46m, Ano46-39, Ano46-84, Ano49n, Ano86i, Ant70, Axt73, Cho58, Dan46, GM09a, Joh59, Lal64j, Neu58a, NAD^{+}47, Opp51b, Pre96, SW10, Ste95, Tas51, Zil64, Ack88m, Ack88n, AS85, Ano90-52, Ano93-35, Ano03i, C^{+}87a, CAN88, CN90, Har79a, HD23, Hec12, Hec20, Mar83b, Mat18, Tsi83d]. Productive [dF79e]. Products [Ano46-58, Ano64a, Axe51g]. Prof. [Ano46-78]. Profession [Ano52g, FU63, Uni57, Ark94e, Fra47, NA91h, Zag95d]. Professional [Ano56c, Cal82, Kat53b, Kat57b, Nad72, Got64]. Professionals [San66a]. Professor [All51, Glu50, Hil53, dF79g, Fie19, Ano70b, BB52, Mer47, Rab50d]. Professors [Fri48b, Wal87b]. profiles [C^{+}87b]. profitable [McK22]. Profiteering [Brz87]. Profits [Ing53, Bro57a]. Profound [Hig76]. Prognosis [Kar99]. Program [A.55, Ano46-28, Ano50-40, Ano51v, Ano52k, Ano54q, Ano93-31, Aug48a, Aur46, Bet93, BGR61, Bon53b, Cra76, DuB63, Fel68d, Foz62c, Gol62a, Gol62b, Gre64b, Haf51, HB10, HPL^{+}49, Hut46b, Hux47, Jac74d, Joh60b, Jon70, Kam71, Lam80, Lap60b, LeR50b, Lin84a, Lun59, MM51, Neu65, NKH03b, NK05d, Pal55, PBG55, Rab72f, Los54, Smi54c, Tru51, Ure67, Ure69b, War48, Whi51b, Wig88, Wu60, AH92f, Ano77j, Arn47, BWL21, Bro87c, Buk97b, DL78, Fel77b, Fra21, Gol79, Gre84a, Hen96, Ham98, Hol77, Kel10, Lyn86, Men92a, NF83, Pfe82, RL76, Roc18, SL80, Tho86b, TM94b, Ano50b, Ano50o, Ano50x, Ano53y, Ano85a, Fox73, Kui73, Nie65, RH65, SS73, Str58]. Programs [A.51, Ano50v, Ano54-31, Ano63y, Bar60, Dan54b, Fru61, LL83, Mor62d, RC70, Ric67, Sim60a, Smi72a, Tur51, Tur52, AK95b, Ano80m, Dun84, Isa80, Kri86, LP85a, O’K84, Rin84a, Roc76, Rot95j, Wan83, YG77a, Zha20]. Progress [Ano50c, Ano50-40, Bor86b, Chu60, Ded62b, Gom68, Hut51, Khr57, Kih47f, Mey49a, Pit52a, Pit52b, Rab46c, Rab47g, Rab68c, Rab69e, Tur51, Alf79, Bar85b, Bor85, Cor78, RH14, Rev21, Sco08b, Ken66, dF79g, Ven80]. Progressive [TR80, Fel79c, Day79, Fel79i, Han79, MSDP80, O’B80, San79, WG82]. Prohibition [Ano57k, Ano88e, Int50, IK21]. Project [Alv87, Ano96-27, Coo76b, Cor56, Ell92, Gra84b, Kor92, Moo99h, Rab68g, Ran53, Smi68, Spi75, Fei81a, Han98a, Joh86, Ne’81, Pai80b, Rot85, AM21, Ano46n, Ano46w, Ano46z, Ano46-51, Ano46-52, Ano47m, Ano51r, Ano86h, Ano96-31, Art71, Boy95, Dra73, Dro23b, FS95, Gou47b, Hut46b, Lew50, Lew63a, Mar66a, McL53, Mec21a, Mon53b, Mye91, Nor02b, Nor03, O’L77, O’N89a, Par66a, Pre63, Rho07, Rho15, Ste60a, SB23, Wid61h, Wig66, Cal82, Gol88b, Moo95d, Pre62, Sea72b]. Projecting [Woo12]. Projection [Sea70b]. projections [Dro24c]. Projects [Ano45i, Gib91b, Luc90b, Sch79a]. Prokofieff [Fel73a]. Proliferation [Ack88f, Act09, AG97, Ano66c, Ano67d, Ano96v, Ano96w, Don89, Fra83, Hal66a, Hal66b, Ing68b, Lal66g, Mab93, Mac97b, Mos81b, Que69, Que70b, Rot84, Spe89b, Bul09d, Ste76a, Sto65b, TK66, Tuc66, WL84, Wei78e, Agn82a, AH91n, AH92l, Alb93d, Alb93e, Ano11d, Cas77b, Dah78, DLF79, Fel78k, Fel79h, Fel80l, Gol88d, GG11a, Hec16, HV18, LM80, Les81, Lev84a, Lev88, Lor77, Lov79, Man77, MW91, Mog81, NA88m, Nye82, Pro85, RS15, San90, San76b, Spe85b, Spe85a, Spe87a, Ste77c, Swe80, WH16, Wei18, WF90, Wil77c, WF16, Yor82, dV78, vHWL79, AC85, Don86b, Eps81, Hal75b, Ima69, Ken63c, Lev85, Cle74, Ste76a, Ano78g, Zim81b]. Proliferation-Resistant [Act09]. Prologue [Ind14, Wil70a]. Prometheus [BS05, Dro23k]. Prominent [Atl47]. Promise [Ano86k, Ano86i, Bro56a, Gar77c, Lam69, LC98b, Mec20f, Mül85, Mec23d, Fin90, Mea93]. Promises [Ano94j, Mit50, Har84a]. promising [Bow89, Hec22]. promoted [Kap89]. promotes [Whe90]. Prompt [Pol09b, Cap98]. Prone [LD00]. Proof [Mak01, Ste07c, Woh78]. Propaganda [Hal93a, Mie76, Sch71a, Ark93b, La 88b]. Proper [You47, Han17, Kul17, Raj17]. Properties [Gro06a]. prophecies [Fre76a]. prophecy [Fel78e, Fel82k, Fel83b]. Prophet [Dow83a, She86b]. Prophets [Aur72a]. Proportion [Ano95p, Ano95v]. Proposal [Ano50-39, Ano51a, Ano54-29, Ano55r, Ano58j, Bar46a, Bet60, Bro61, Chu60, DB50, Eis54c, GA73, Gne77, Gro46, Gro06b, Hil60, JC50, Lal64j, Led66b, Ler46, Mar48e, McC57, Rab55a, RE68, Sen65a, Sto72, Str54a, ZSA73d, Ack88b, Cop80, DPT17, Dje76, Gay83, SF81b, Yaz88, vMK88]. Proposals [Ano46b, Ano47d, Ano48-40, Ano52-27, Ano55y, Ano60j, Ano66l, Ano73h, Ano77m, Ber84a, FGJ65, Fro49, Gre67a, Gro47c, Hil65b, Mil50, Pel64, SD56, Tay56, Voh71c, Ano87-28, Bro86b, Lan89d, Lin79, You66a]. propose [DS47b]. Proposed [Ano48b, CH54a, NA92i, RS88b, Ano92-31]. Proposes [Rob49, Gar83b, SG83]. Propositions [Mer47]. Propper [Eva86]. Propulsion [Ano49-41, Coo62b]. Pros [Pra04c]. Proscription [Wes74a]. Prosecution [Gla75a]. proselytize [Ano93s]. Prospect [Sci49]. Prospective [Cou54]. Prospects [Ano71e, Bas80, Cha66, Day75h, Fly97, Git69, Gro53b, Hau64, HGV10, Ing54b, Ing57b, IS67, Joh59, Mea85, Mei51a, Mei51c, Opp55, Por51, Pry48, Put52, Rab69e, Sch78a, Smi69b, Sne57, Spo50, Spo53, Top62, Vin59, Wei58, Wie70, Wou81, You65b, Bar77c, Isa98f, Pan88, PW76a, Mul66]. Prospectus [Spa69]. Prospering [Ano91-51]. Prosperity [Gor95d, Hol75, Pea58, Bus94, Ano78o]. Protect [Bli70, Mor00, Ure48d, Cot22, Gou24, Len20, YWS20]. protected [BW84, Nov77b, WB85a]. Protecting [HMR86, Sca91, SLM87]. Protection [Aft99, Ano50-41, Ano78g, Ano78n, DM05, Fre76c, Gar83a, Jac74d, KMN04, MK02b, Voh71a, Yar80a, Ano95l, Ano97n, Eps77c, Meh95, Row77, vH80, Gue80]. Protective [Coa47]. protects [Rot18]. Protein [All62a]. Protest [AAF^{+}52, Ano50t, Bak75, BS47, Bou65, BCH^{+}48, Eps88a, Gen71, KASN52, Kur79, Swe86, Swe89a, Tis52, Wit04, Wou76, Ad 90, Ano78i, Ano82c, Cha78a, Gor80, MJ81, Ray80, RS88a, Nan73]. Protested [Hah61]. Protests [DMD64, Pol53, Wil15]. Protocol [PR08, Bel72, Mes72]. Proud [Zie77, Ano92w]. Proven [Fly01c]. Provide [Hum55]. provides [Hol77]. Proving [Ama77, Fed70, Har77b, Sta00, Str56]. Provisions [Ano46-44, Ano50w, Oom46b, Smi93a]. Provocations [Ark85g, Ack88g, NA88l]. Provocative [Gar97a, Coe98]. Provoking [McK89]. Proxmire [Rod70]. proxy [BR79]. Prussia [Dow83f, Ano83b]. Prussianism [Smi67]. prying [Gri91]. Pseudo [SB77, Shi76]. Pseudo-McCarthyism [SB77]. Pseudo-Scientific [Shi76]. PSR [Ano91-32, Ano91-52, Cal86b, Rot93w]. Psyche [Hal07]. Psychedelic [Fre68a]. Psychiatric [Bet65]. Psychiatry [Ano52f, Lon52a]. Psychic [She05b]. Psychicalistic [Har64b]. Psychochemicals [FL62, Lie62b]. Psychological [Car82a, Cha53, Fra76b, Jan50, Keu83, Per91a, GCV16, Wel16, vH11, Ano55c]. Psychologist [Jor63, McN63c]. Psychology [Ano52f, Kow49, Lon52a, Mac61, Mar54c, Mer54b, Saw63, SJ64, Sti60, Wyl54, ST22, Tre18, Ano87p]. Psychopathology [Mic62, SK69]. Psychosocial [Mac81b]. Ptolemaic [Sha54b]. Pu [Moo97n, Rot93a, Rot96j]. Public [AP80, Alp62, And65, Ano60f, Ano76d, Ano78g, Ano84h, Bak75, Bal67, Fal68, Gom70, Har79b, Hei73, Hil53, Hir76, How56, Jas81, Kal68, Kel64, Kil59, Lan89e, Lau46, Law88, Lee74, Lev60, Lev53, Lil48, Lil49b, Mar48b, McK74, Mea69, Opp51c, OS00, Pre48, Pal64b, Par46a, Pas77, Pei69, Pom61, Rab57h, Rab62b, Rad53, Rie59, Roc83a, Rot79a, Rot03c, Sci55a, She64c, Sig85b, Sim47, Sko68, Bul08a, Ste62, Ste75b, Suz84, Ter58, Tuc72, Tur53, Wei68a, Whe84b, Zin79, Ahe90, Ano95m, Ano98l, Ano15b, Bar81h, Bro86a, CYP14, Cor94, DM24b, Dow78, Flo76b, FWCL20, Ger21, Gre78, KW15, Kal82i, Khr10, Kul95, Lak78, NR79, NK15, Nov77b, Pai80b, PD12, Ram11, Rep83, Ros78a, Ros89a, Ros90a, Smi82b, Sol79, Mar72a, Sno69]. public-interest [NR79]. Publication [Mey61, MBD^{+}62, Rab48a, Rab46d, Wei46, Woo62, Lew72b]. Publications [Ano46-79, Gun66, HB69, For80]. Publicity [Shi56b]. publish [STRS80, TR80]. Published [Ano46v, Nor04]. publisher [Ano75e]. Publishes [Ano46d, Ano46t]. publishing [Ano98k, Fie21a]. Puchala [Sco74]. Pugwash [A.62, Ano58g, Ano58l, Ano58k, Ano59d, Ano61i, Ano63o, Ano70a, Ano79n, Bro58, Fed59b, Fed59c, Fel81c, GH85, Mah59b, Mah59a, Rab67h, Rab70c, Top58b, Wei58, Ano57i, Ano60t, Ano62x, Ano64o, Ano64p, Ano64q, Ano66p, Ano67l, Ano71b, Ano74i, Ano79d, Ano80g, Ano81j, Ano85g, Ano85x, Ano86-36, Ano87n, Ano87-33, Ano88q, BGR60, Can79, Cah71, CW81, Com57a, Com57b, Com57c, Fel73d, Fel78l, Fel82i, Gan83d, Hol84a, Int67, Kal82g, Kap80, Kap83c, Kap89, Kra57, Lan90n, Mar58b, Mar58c, Moo97i, Moo97k, Pai82j, Rab57c, Rab57g, Rab58f, Rab60e, Rab63e, Rab65a, Rab65d, Rab67h, Rus58a, Rus78, Smi97, Thi58, Top58a, Vin59, Ano78l]. Pulitzer [Dro17]. Pulitzer-winning [Dro17]. pull [Ste85b]. pulse [Ste83a]. Pump [Cav06, Sea85b]. pump-priming [Sea85b]. Pumping [Sch04c]. punch [Lym11, Rot94d]. Punishment [Str06c]. Purchasing [Loc94]. Pure [Ale69, Coc52, Eme65, RB65, SW59, Sno68b, Sto76b, Ure55, Wei65, dH75]. Purge [Bri50, Dun49, Rab49e]. Purification [rWeKHF47]. Puritan [Jac83]. Purity [Lic77]. Purloined [Rot01g]. Purpose [Cla68, Day74a, Men53, Gro81]. Purse [NA91m]. Pursuing [Ano11h]. Pursuit [Hor75, Sai07, McN82]. Pursuits [NK03b, McD13, NA88i, NA89k, NA91j, NA93k, NK12b]. push [Gho15a, Kas21, Kre97, Ste85b, WF86]. Pushes [Ros75]. Pushing [Ano93x]. pussycat [NA89o]. Put [Day75f, Mat84, Per91b, Zam92c, Isa92c, Isa93o, Pat78b]. Putin [Dro22h, Hil16, Hil20, ST22, Zub22]. Putnam [Kat54f]. Puts [Ano46-50, Eav17c]. Putting [Ano94o, Los57, Pan84, Vog13, ZSA73d]. Puzzle [Bre69b, BR97, AH92h, Bey12a]. puzzlement [dTM93]. puzzles [Hib94b]. Puzzling [Fra21]. Pyenson [Ber79]. Pyongyang [Ano02k, Lee24]. Pyotr [The90, Fel84f]. Pyramid [Ams76]. pyramids [Rot89a]. pyrrhic [Ano93s, Isa85b].

Q [Gre64b, Han95b]. Q&A [Dro20b, Dro22i]. Q-Clearance [Gre64b]. Qaddafi [Pra05c]. Qaeda [Nas03b]. Quad [Ano95l]. quagmire [Bar10b, HI94, Kib10, Kor10, Lan10, Pic10, Ram10]. quake [NA89a]. Quaker [Exe57, Edg52]. qualified [Eps90b]. Qualitative [Yor73]. Quality [Bar58b, Cah75, GF73, Gom70, Mac65b, Nor74, Pec70, PF75, Sin81]. Quandary [Raj10, Jas24, Kal83b, Pat81c]. quantitatively [Eps88a]. Quantity [Bar58b, Cah75, Nor74]. Quantum [Adl69]. Quareshi [dF79g]. quarked [NB78]. Quarks [Ano78m, Fla90]. Quarter [Shr70]. Qubain [Ker68]. Quebec [Ano58g]. Queen [Lev73]. Query [Fla80a, Alp92, Ano97g, Mar82b, Per97]. Quest [Ano57k, Gla56a, Hoe09, Inf06, Per00, Rin61, Kar88c, Kor16, Tod81, Dan57]. Quester [Cle74]. Question [Aur71f, CR99, Dav48, Dun75, Fin65, Gol48, Hen71b, Ing53, Kri75, Moo00f, Roc81, Sal77, She60, Sta61, Ste80a, Don88, Mar85b, Sic96b, ZRG83]. questioned [Lei90]. Questioning [Hib95c]. Questionnaire [Gro47a]. Questions [Adl57, Hec16, Lal64j, Pal00, Sim59f, Zim76, AH89a, Day75a, Day76h, Mec16e, Moo94f, RL78a, Sok16, Wio77, Moo00a]. quibble [LH91]. Quick [Wyl56, BK81, KB80, Leb87, MM90, PWS^{+}96, vHWF^{+}81]. quickens [Spe85a]. Quid [Rab59d]. Quiet [Ben01d, Ehr72, Jac63a, Lew64b, New62b, Puz93b, Ano93v, Ano98m, Ell77, Nel90, NA92e, Zam92a, Pal64b]. quietly [NA93l]. Quills [Ber76a]. Quincy [Ano46-80]. Quite [Hil51]. Quixote [Cha78b]. Quo [Law69, PG84, Man85, McL11, Pac59, Rab59d]. Quonset [Rot89a]. quota [Dre77]. Quotas [Ano50q]. Quotes [Ano51x, Ano52u, MCC51, WB51].

R [Ano60f, Ano78n, Ano81f, Ano91s, Ano11e, Bal79b, Bei84, Ber03, Bri80b, Bro47c, Cal62, Che88, Cra83a, Das96, Eck78, Ehr81, Ehr83, Fei92, Fel82g, Gen86, Gol88b, Gol78a, Han75, Har92b, Hob85, Hun87, Jas81, Kat52b, Klo85, LaF86, Lan57, Lap55a, Leb91a, Mye89, Nei82a, Nor02b, Nor03, Per95, Pla57b, Rob69, Roc85, Rub73, Saw80, Sch86, Sco74, Sha50, Sha81a, Shu78, Shu49, Smi85d, Smi88, Som91, Soy89, Bul03a, Bul08o, Bul08s, Ste76a, Sto58, Tod79, Wou81, BN67, Hol64, Isa81a, Mar64f, Nel64]. R&D [Abb90, Din78, Gol91a, Gol93a, Her78b, Kyo85, Kas69, MS77a, MS77b, Nie66a, O’L75, RL76, RL78a, Rep85b, Tsi77c, Tsi90, Yar80b, Yar80c, GG11a, Dow81a]. R2D2 [NA92a]. Rabbitt [Sca91]. Rabi [Ano88m]. Rabinowitch [Ada97, Ber81, Cho81, Kal68, MDB81, Shu81, Ste81a, Ano72f, Ano73c, Ano73d, Ano81m, Ano82r, Elm74, Gra82b, KFM50, Mye82a]. Race [And56, Ano54t, Ano54u, Ano54v, Ano74c, Ano87i, Ano87-39, Ark87a, Ark87b, Asi62a, Bar87, Bat46a, Bat46b, Ben02, Cor69, Dee57, Dow80e, Eps77a, Fen91, Fer60, Fie87, Gar67, Gar81, Gou46, Gra73, Gus83, HD82, Hal75b, Han87a, Ing56a, Isa05a, Jac74b, Lei72b, Lei74, Lew70f, Lew70g, Mcc87, NA91a, Nor64, Pan71, Pos87, Rep85a, Rui68, Sim57b, Sto64, Sto65a, Str72, Wie67, Yor69, Yor70, Yor73, You66a, You66c, Ano11f, Boy89a, Bre82, Day77a, DeW84, Dow78, Dow82h, Fal87, Fel84c, Fra76b, Gal81a, Gei16a, Hes85, Hol85b, Kis79b, Mac81b, McC84, McD85, McM05, Mor83c, NA91c, Pie80, Ric79, Rof19, Sco77, Smi85a, SS88, Ste77b, Uni78, Zal19, Ano90p, Bet69a, Cow65, Deu83b, Dow80e, Dow83f, Jen70, Ano91s, Cra84d, Don89, H.84b, Rob91b, Tsi72a, Tsi83c]. Races [Ise92, LC09, Kla97a]. Rachel [Ehr79, Moo95d]. Racing [Nor02b, Nor03, Rao87]. rack [Rot95g]. Rad [Ano91-53, Ano98m]. Radar [Gar91, Swe87a, Bur87, Eav17c]. radars [Rub87]. Radiant [Gen86]. Radiation [Ato54b, Ahe90, Alb71, Alv80, Alv84, Alv87, Ano56n, Ano63x, Ano78g, Ano91-42, Bac58, BS12, Bru48, Bru58, Bur47, Cha58a, Cla56, Com63a, Com75b, Con90b, Cro58, Eri91, Err58, Fin90, FJ12, Fra58, Gal90b, Ham72, Hol58, Hol56, Kuz58, LS58, LeR50b, Lib58, LR71, Lin85, Lis46a, Lor01d, Lou90, McC90, Mei90, Mor59b, Mul55b, Mul55a, Neu58a, Neu58b, Oga58, OW90a, Osb58, Pau70, Rot93w, Rum71b, Rus56b, SSW^{+}55, SL58, SE58, Slo12, Ste90a, Ste72a, Ste90b, TG70, Tam71c, Ter58, Uni57, You78c, You58, Abr86a, Agn77b, Ano77n, Ano86-33, Ano95i, Ark97c, Bey12b, Bey12a, CMR79, Coh79, DM18, Fri11, Gro82, Har91a, HQL84, Kas12, Lof84, MK80, Mie77, Mor78, Nus85, Rot78b, Rot81a, Sal90c, Sol79, Tho12, McC96, Rot79a, A.54, HI72, Kat54e]. Radiation [A.54, Ano53e, Gar83a, Lis58, Rad83, Rei68b, Upt63, Wil92a]. radiation-free [Ano95i]. Radiation-Induced [Wil92a]. Radiations [Gla56b, Lac57]. radical [Gar88d, Kok88b, Sig89]. Radio [Ano45c, Ano93x, Fos88, Joh94c]. Radioactive [Ass80, Abb79, Ano46n, Ano48n, Ano56c, Bar90c, DeN74b, Dow80j, Gen59, Hol58, Lap55d, Lap55e, Law84a, Lew71a, Lib55, Rid50b, Spe74, Ter58, WZSA73, Woo73a, ZSA73d, Abb84b, DM24b, EG24, JG96, JS79b, Moo14, MM11, Pen84, Sal90a, Ano71g, Ano78n, Met83]. Radioactivity [Lap59f, Sal90b]. Radiobiology [Kuz58, Led71b]. Radiocarbon [Lib48, Mat56]. radiofrequency [Sol79]. radioiodines [SR81b]. Radioisotope [Ano49-33]. Radioisotopes [Ano47-29, Uni47, She14]. Radiological [Ano46q, Ber85, Kli76, Kup51, BBP14, BB20, vH11]. Radionuclide [ZSA73d]. radiophobia [Rop16]. Radon [Coh76a, VR67, Coh76d]. Radwaste [Met83]. Rae [dF79e]. Rafferty [Dow83e]. Raft [Hey72]. raging [Isa92a]. Rahman [M.84, dF79g]. Rahn [Ano85d]. raid [WK82]. Rain [Ano69h, Col84c, RC90, Kal82i]. Rainbow [Say84, Ano91-54, Dow80e]. Rainforest [Ano93m, Sea01, Dro25a]. Rainforests [NN80]. Rainmaking [Mac68a, Ney69]. Rainwater [Sch68e]. Raises [Fav52, NA91g]. Raison [Pug73]. Rajasthan [Her83b]. Rally [Ano46-45]. rallying [Wil90]. Ralph [Ano87p, Dow82d, Fin57b, Gre58, Gre64a, Kal63a, Lal70, Lis58, Mar58a, Pre62, Rab53a]. Ralston [Ano86k]. Ramage [dF79h]. Ramakantha [Dow79b]. Ramberg [Coo81]. Ramet [Con88]. Ramifications [Spi65]. Ramo [Sei81]. ramp [Doy22]. rampage [Pop93]. Ramsey [Mar54f, Tho62]. ran [Sto19c]. Ranch [Pfe82]. Rancho [Knu89]. Randall [Ste95, Ano83f, Mec21c]. Randolph [Wyl56]. Random [Fel76f, Tam74]. Range [Hol73, Rab50a, Smi73, Szi60a, Tsi75b, Wis55, NA88o, WW60]. Ranum [Tod81]. Rape [Kre98]. Rapid [Ano94o, vHC76]. Rapid-fire [Ano94o]. rapidly [Sie24]. Rapoport [Bre65]. rapped [Wei94e]. Rarotonga [Sut96b]. Raskin [Mul65]. Rasmussen [vH77b]. Raszak [Hym68]. Rate [Opp51b, Str67, Abr83, GM90]. rates [Rot77]. rather [TS22]. Rathjens [Tsi72a]. ratification [Isa88d, Kas21, KT97, Mat83a, Pif24, Pod24]. Ratify [Nel09, Mei98, Men91c]. Rating [Swe80]. Rational [Leg57, Osg62]. rationale [Ano91-36]. Rationales [Kra07, McM65]. Rationality [Aro64, Mic61, Phi94]. rationalize [Ano93u]. Rationalizing [Moo97a]. Rationing [Smy51]. Rattinger [Sig85b]. rattling [Wei78b]. Raul [Ano89e, Ano87n, Men93b]. Raviv [Mos91]. Raw [Sch62b]. Ray [Tel46b, Woo78, Ano86-32, Blu88, DeW88, RK91]. Raymond [Fre88, Gal86, Mes83a, Sei62, She61, Smi58f, Uhl95a]. Rays [Uni57, Sal90a, Ano78m]. razor [Ano95m]. RBMK [Bla87b]. Re [Ano74b, Coh48, Day74e, Eme47, Ken71, LD79, Mul65, Mul73a, Ore62a, Sim54b, Woo00e]. Re-Appraisal [Mul65]. Re-Define [Eme47]. Re-employing [Ken71]. Re-Examination [Coh48]. Re-Reading [Mul73a]. Re-Set [Day74e]. Re-Trial [Sim54b]. Re-Up [Woo00e]. Re-Use [LD79]. Reach [Coo06, Coh86, Wal21b]. reached [Kam20]. Reaching [JFS20, Rot98e, Wei71]. Reaction [And73, Ano54g, Ano54y, Att54, Chu54a, Dea53a, H.85, Hoc57, Kra57, Rab57c, Rab05b, Sel69b, Sko58, Wat74, Bro79a, DuB62b, Eis62, Fre77b, KS98]. Reactions [Ano62y, Ano69l, Ano69k, Cun52, Dav57c, Fel69b, GBG^{+}81, Han69, Kis69, Kur56, Pla61, Rab69a, Tel56, CN88, TFM81]. Reactor [Abr74, Ano49u, Ano49-41, Ano50-28, Ano50-40, Ano52k, Ano54q, Ano62-30, Ano86u, Aur72b, Bea87, Bet75a, Bet75b, Cha75b, Gol90a, Haf50, Haf51, Lov73, Mei61, Men53, O’C87, PvH74, Pri75b, Pri75c, Sco90, Smi74a, WW61, vH76a, Ano91-59, Bey80b, BvH82, Bro86d, Eav17a, Fre77b, Fly99a, Her83b, LaM79, Lon52b, Moo96l, Ner80, Pat86a, Rad97, Ram86a, Rei91, Sai10, Ste89c, Sto90, Tak17, Tho80h, WM78, Zin53, vH76b, Hyd91, Ste89d, Tho86a]. Reactors [Ano71d, Bac49, CFM^{+}10, Den53, Fei77, FS08, Fer70, Gol09, Jac74b, Lit78, Spi60, Tor77b, Wei59a, APvH86, Ano89u, Bra17, Coc81a, Coo12, CJLT24, Dav79c, Hal93b, HMR86, Hoe87, Hor88a, Hor88b, Ikl80, Jan15, Jas24, KM18, Loc14, Mab93, MR21, Moo15, Nat85b, PR14, RR15, Ram16, RS88b, Rot77, Ste86b, The88, Tho84c, Zha15a]. Reacts [Khr99b, Ade82]. Read [Ano54z, Bar57, Ano86v, Ano91-39, Ano98l, Flo88, Bro61]. Reader [Bor58c, BKSG60, Dee57, Dun74, Fed61, Fel61, GLMS60, Gun66, HB69, Lap59g, Mar59b, MS61, McC56, Pin57, Rot58b, Ste60b, Wei61, Wer56, Wu60, Ark95e, Div79, Sch68f]. Readers [Ano53-44, CMMF62, HJS^{+}92, RG49, LaB81, PGL91]. Readiness [Gor95e]. Reading [AM51, Bol88b, Dvo07, GS96, MR09, Mul73a, Sig10, Spe75, Tuc10a, Ano78k, Ano78n, Ano78l, Ano78o, Ano78m, Ano78p, S.80, Tho80g, dF79f, dF79h, dF79c, dF79e, dF79b, dF79a, dF79g, dF79d, dF80]. Readings [You66a]. Ready [Bla07, Gro05, Han96a, Hav03, HW08, Isa00d, Isa02e, NA90l, Wer56, Ano92w, San91, Smi92c]. ready-to-leave [Ano92w]. Reaffirms [Ano48-37]. Reagan [Dia83, Fre84a, Rau91, Swe88, Wei89a, Ada83, Ano87-34, CA87, Can70, Cla81, Ger82, Gre83a, Har84b, Har87e, Has85, HC87, Isa81b, Isa81a, Isa85b, Isa86c, Isa86d, Isa88b, Isa01a, Jac82c, Kin81b, Kre86d, Kre90, Lan82b, LD89, Lin86, Men86, NA88n, Rot95f, Sig88a, Sig89, Smi87c, Som84a, The89f, Uhl95a, Cra84e, Gal86, Mes84b]. Real [Ald17, Ano05b, Aud65, Aue06c, Boe76, Bro56b, CPV01, Dow81f, Fau87, For09, Gyo77, HWH78, Hur78b, Kih47f, Lin84a, McM52, Mil03, Pra03c, Rab51g, Sta61, Sut77, Tri71, Ure61, Ano93-34, Ano97m, DE95, Ema21, FR98, Gap90, Goo04a, Goo04b, Gor96a, Hib95a, Moo95d, O’S86, Ode88, Smi77b, Sut16, Wol91b]. realignment [Den14]. Realism [Aro55, Bob65, Kas67, dSP60, Ano93y, Zag92]. Realistic [Iri56, Lon57, DL21, Dro24c]. realistically [Mec19b]. Realities [Bla85a, Der66, Der69, Gan83b, Har65, Sta73, BKV13, Ikl76, Osg83, Sal76, Sim54a]. Reality [Ano96j, Bor58a, Bra09, Ede89, Gre63a, Har77a, Pra04d, Sco94a, Spi71a, Bul01c, Wal64, And81, AC86, Cas78b, Eps87b, Far98, Kin81b, Mec21b, Men91a, Mil12a, NGG11, Rot96k, Sep23, Ehr82b]. Really [Ano94o, AM74, Bro64, Gil69b, HM90a, Ich64, Ing55a, Mor56, Myr71, O’B69a, Pai03, Pal01, Rot02a, Ano92r, HSV19, Kla19, Sal88, Mye95, Dan52]. Rear [LaF90a]. Rear-view [LaF90a]. Reardon [Ano88e, Kre90]. Rearmament [Bar58a, DM81a]. Rearranging [Joh95]. Reason [Bal60a, Con46a, Fox61, Ing74, Ano94m, BM99, But98, Kle81]. Reasonable [Rep86]. reasoning [Mol89]. Reasons [Moo93i, Nel09, Rob08, Rus62b]. reassert [Tru85]. Reassessing [Nel10]. reassessment [Nad80, Tod79]. reassessments [Ack89d]. Reassurance [Day74d]. reassure [Wu23]. reassuring [HSV19]. Reawakening [Mul64b]. Rebecca [Smi58e]. Rebellion [Abr73, McG68, Myr72, Rab68e]. Rebellious [Wie48]. Rebels [Wie47a, Mor85a, Olu71]. rebirth [Abb90, Moo94c, Tsi90]. Rebuilding [Ada87, RE19, Ano93m]. Rebuilt [Rab51k]. Rebuttal [Coh76d, Dic77, Lap76, Roc76, Sch75a, Sel77, Zel76a]. Recall [Wig61, NA90l]. Recalling [Ada97]. recalls [Ano90s]. Recasting [AG86]. Received [Ano49i, Ano50e, Ano50f, Ano51c, Ano52d, Ano53f, Ano53g, Ano53h, Ano53i, Ano54d, Ano54e, Ano68a, Ano72c, Aue05b, Ken61b, Khr61, YTS61, Ano50d]. Receives [Ano52-28, Ano85u]. receiving [Opp54a]. recession [Spe78]. Reckless [Mas57]. reclaim [Lif13]. Reclaiming [Sch90]. Recognition [Wri65, Leo19]. Recognized [Bus47, Ano93-29, Sch21]. recognizes [Ano89p]. Recognizing [Fel78f, Pac59]. Recollection [Sea70b]. Recollections [Bor65a, Bor65b, Bor65c, Gro70, SW80a]. Recombinant [Bal78, CT78, Day77g, Fed78, Nov77a, Tho79c, Tho80g, Wri77, WS83, Gre78, Kri78, WSH^{+}84]. recommendation [Gle21]. Recommendations [BBF^{+}82, GNMS54, Gor51b, McE52, MB48, Ano96x, DG07, Int86b]. Recommended [Spe75]. Recommends [Ano46-88]. reconciliation [Kat80b]. Reconciling [Bar80h, dF79f]. Reconnaissance [YG77b]. Reconsideration [Bar67c]. Reconsidering [Ber03, Ric84]. Reconstruction [Rei78]. Record [Ell93, Mor61b, SH61, Ano97t, Bal89b, Ber86a, Con90a, Div86, Isa88b, Mas83, San90, Lin84a]. Records [Ano45c, Ano91-53]. Recount [SN89, Ano97l]. Recruit [Wil75f]. recruits [Ano92t]. Recycle [MC75, STC74, Dah77]. Recycled [Ano91-42, Ano93y].

Recycling [Ano91-55, Ano97m, Ber72c, Isa92i, Kra73, Day76c, Zha20]. Red [Ano92v, Ark87h, Cre63, Dow81f, Dow81h, Khr96b, Rab65c, Ros90a, SC76, Sim87a, Smi67, Ano91-53, Ano91-59, Bin79, Dor85b, Gol93c, Pen98, RK93, Uhl98b, Ano52j, Ano80o, Ano92p, Arb93, Int50, Sch89c, Lee91a]. red-tinted [Dor85b]. Rededication [Sto59b]. redeeming [Ano93y]. redefines [Duf86]. Redefining [LBCA23, Lug05, Nag84, Som91, Sør81]. Redefinition [Mul73b]. Redevivus [Rab59h]. Redick [Sha81a]. Redirecting [Kan09]. Rediscovering [Mye93a]. Redressing [Kaz14, Mal14, Nie14, Sin14]. Reds [Gre64a, Day77b]. Reduce [Lar65, Coa47, Fie19, Fie21a, Kor21, SMĆ08, dTdLF24]. reduces [Sez20]. Reducing [Coa47, FUvH85, Gol93a, GM09b, Sto18a, Boo22, WH16, vD78]. Reductio [She64b]. Reduction [Ano47q, BGR61, Chi59c, Etz63, Sil64a, Ano84n, Bac81, Ble88, Ble89, Bor87, Knu89, Lan89g, Smi16]. Reductions [Ano78o, Ark99c, Lal64f, Lal65g, NK04d, Rot91b, Hin18, KN14d, LE85]. Redux [Sco08a, Ken87a, Leb91b, Rot94k, Sco08b, NA93j]. Redwing [NA92f]. Reed [Mec24, Mil85c]. reefs [Ano94j]. Reel [Dow80e]. Rees [Bro20]. reexamined [Cap98]. Reexamining [Mar63]. Referendum [Bar80l, Bar79c, BL83b, NW79, Pat79b]. Refined [Sma76b]. Reflection [Sea79, Kap79]. Reflections [Ano91-67, Ber72b, Bou65, Cah94, Con49, Dys61, Fel84b, Fre68b, Gow79, Gre73a, Gué66, Han66, Har79c, Hoa59, Hol76b, Jac74c, Kea61, Ken61c, Khr10, Kim87, Mal54, Mar79c, Mar61c, Mor70b, Mul73a, Pei85, Pet61, Rab69h, Rot61, Roz64, Sei80, SW10, Sto61, SG69b, Til64, Tod81, Wei78a, Wie56, Wil67a, TR80, Tur77, Bor65c, Dow80i, Duf70, Fla68, Mot68]. reflects [Haz82, Hoa77, McN82, Wei86a]. Refocusing [Gar87]. Reforging [Pan86a]. Reform [Ano75c, Ben87, Cra84e, Mil88, Mil85c, Pra96, Joh94d, RL76, Mor87c]. Reforming [Gil72, Har14c, Ros67, Rus92b, Fie21b]. reforms [Hof88]. refuge [Ano91-39]. Refugees [Lar07, Loe02]. Refusal [Wil62]. refuse [Ger96, Sol80c]. Refusnik [Ses83]. Regarding [Ano60a, GLMS60, Lap59a]. Regents [Law84b]. regime [Cra87, Mec20b]. Regimes [Bal79c]. Regional [Ack07, CR71, Hun87, IW52, Jon90b, Kin61, Sub84, Dea91, RG13, Van22]. Register [Ano78o]. Registering [Gor80]. Regression [Bro57e]. regrets [Rom91]. regulate [Cha16, Rof16, San16, Wri77]. Regulating [Kri14, O’B79, Per87b, Tho79c]. Regulation [Ano47q, Gal65, Ahe91, BS12, CZ95, Gre12, Vic78, Ano78n]. Regulations [Ano48b, BFW47, Gil55, Gre68a, Wat77]. regulators [Gil18]. Regulatory [Jas81, Scl79, Sto19c]. Rehearsing [Gol88a]. reheats [Lan90j]. Reich [vH77a, Hei68, Rab71b, Gou78b]. Reiffel [Kin80c]. Reign [Sil79a]. reignite [BHK^{+}21]. Reimaging [Ger21]. Reinhardt [Lap55a]. Reining [Mol03]. Reins [Gol92c]. Reinterpretation [Fre88]. Reinventing [Gol93d]. Reischauer [Lei68, Sol68]. Reiser [Sil79a]. rejection [JC54]. Rejects [Ano48-40]. Rejoinder [HE72, Mel53a, Ney69, Ada86b, Har85c, Kes85b, Leb88a]. Relabeled [Moo02c]. related [Gib91b, LST10]. Relating [dRMT^{+}47]. Relation [MEA^{+}47]. Relations [Adv50, Cra85, Dub55a, Fox48, Gal86, Gus83, Jor63, Kal70, Kra46, Lal67f, Lal67h, McN63c, Pol09a, Bul08c, Tuc66, Uhl95a, Win90, Woo56, Ano84e, Bal85, Blu84b, Brz82, Car84a, CS16, CG12, Cra84a, Fel77o, Joh93, Ken78, LD89, Mar82a, Osn77, Sak84, Sal16, Sez20, Shu87, Wu22, vL85, Tho84a, Bro57b, Cav57]. Relationship [Ahn10, ABKS10, Dub56a, Ken81, NA90n, Zha19, Lei92]. Relatively [Cro07, Mat12]. Relativity [Cha75a, For80]. Relax [Fle58]. Release [Hug56, Lil62, SR81b]. Released [Ano48-38]. releases [Ano90-43]. Releasing [Ano91q]. Relevance [Bur65, Fla68, Kan71a, KP71, Lew70a, CK20]. Relevancy [Aur71f]. relevant [Kam23]. reliability [DM84, Lar84, McK86]. Reliable [Dre07, O’S63, Sch01a, DL21]. Reliance [Kli81, Ano78l]. relic [Tho90]. Relief [Rei78, Tak84]. Religion [Bel57, Bix49, Cla59, Con88, Dub56a, Gen84b, Mor57b, Sac57, Sin56, Vis56, Vis59, Woo98e, CO14, Lan76, Pre76, Mor59a, Rap65a]. Religious [Ano46-59, Hes85, Lin84d, Luc82, S.80, Ano15g, Fly98a, Lon50]. relocation [Lam83]. Reluctant [Mti96b]. rely [GM87]. remain [Mec19h, Ram13b]. remains [Mei20]. remake [NH08b]. Remaking [Fra83]. Remapping [Gol92a]. Remarks [Lew70f, McN67a]. Remedies [Har54, Spe54]. Remember [Ano94o, Arp61, Day76d, Rot96a, Smi81, Woo01a]. remembered [Mag81]. Remembering [Coo04, GDH^{+}04, Moo00g, WRT93, Mec20g]. Remembers [Leb91b, Kap90a]. remilitarization [Isa91q]. Reminder [Rin75]. reminiscence [Kis80]. Reminiscences [Sue68, Wil74a, Wil79, Qua80, Wil75d, Fel81a]. Remnant [RHKK78]. Remote [KKS22]. removal [Sto19a]. Removing [She86a, All54b]. Renaissance [Bra09, Bro72, Nor64, Nel10, LSB74]. Rene [Rus62b, Don86a, Gan84, Hul76, Tod81]. Renegade [Lee91b]. renew [Wil90]. Renewable [Ano89l, For09, Mac22, Smi79, DL21, Dro22e, Gho15a, Mye91, Pal16, PD12, SS80, Pat81a]. renewables [Dro16a, Dro21a, Hec22, Mec16d, Mec21b]. Renewed [Nat74, Isa94b]. renewing [Sal80c]. Rent [Nan73, Ano97n, Kre89b]. rent-a-spies [Kre89b]. renunciation [Int78]. Renzo [Dow83f]. reopen [GP92b]. Reorganization [Ano46-94, Lon73]. Repercussions [ABB^{+}46, Ano46-81]. Replace [DeC48, Wei66b, Wes79a]. replaced [Gre12]. Replies [Ada47, Bro69, Bus47, Doo60, Dys65, Ein47, Ell47, Fer61, Gro47a, Har47, Hux47, Hym69b, Kan71b, Led67, Lev65, Mar61c, Mer47, Mor47b, Opp54b, Opp23, Pol62a, Rus71, TWSM61, WC47, Wie47b, You47, CB95, Ein63, Kra89]. Reply [Ash50, Bla69b, Bro51, Ein48c, Ein53, Fri48b, Gro47b, Jen70, Led72, Lil49a, LM61a, Mor48a, Rab59c, Sch72b, SD55, Spi72b, Ste69c, Ste69d, Ste70, Tri72, WKW57]. Report [Adv50, AEC49, AJSB79, Ack94, Alb81, AH91m, AH91n, AH92k, AH92l, AH93a, Alb93e, Alb93f, AG97, ABB^{+}46, Ano46f, Ano46-38, Ano46-54, Ano46-81, Ano46-89, Ano46-88, Ano46-95, Ano46-99, Ano47j, Ano47s, Ano48g, Ano48-28, Ano48-29, Ano48-36, Ano48-39, Ano49d, Ano49f, Ano49l, Ano49x, Ano50u, Ano50-29, Ano50-36, Ano50-48, Ano51l, Ano51y, Ano52a, Ano53d, Ano56j, Ano58i, Ano60c, Ano60d, Ano61r, Ano63y, Ano64a, Ano67m, Ano68j, Ano71f, Ano72h, Ano74i, Ano80q, Ano83d, Ark93b, Arn90, Arn91, Arn96a, Arn96b, Aur72f, Boa91, Bio73, Bab93, Bai92, Bar98, Bar79a, BAB^{+}46, BW94, BEGJ72, Ber72d, Bet98, Bet75b, Bin79, Bos51, Bru90, But97b, Cai90, Cal61a, Car49a, Car94b, Car94c, Cav92, Cla91, Cle92, Coc51, Com73, CCMU46, CF46, Cra60, Dan91, De 70b]. Report [Dow81d, DCM87, DW90b, Eli46, Eps91, Eps92b, Eps97, Eva54, Fed55, Fei92, Fel80l, Flo90, FHN^{+}46, FHN^{+}63, Fre63, Gal91, Gan72, Gar68, Gar91, Gar81, Gar98, Gen72, Gol90c, Gol90d, Gol90e, Gol91c, Gol94, GPR81, GM54, Gra94, Gro53c, Han75, Han91, Han93, Han96a, Hib95e, Hib95f, HAIT80, Int86a, Iln97, IS67, Isa90z, Isa91r, Isa92n, Isa93o, Isa94f, Isa94g, Isa95a, Isa95b, Isa95c, Isa95d, Isa96c, Isa96d, Isa96e, Isa96f, Isa97a, Isa97b, Isa97c, Isa97d, Isa97e, Isa97f, Isa98b, Isa98c, Isa98d, Isa98e, Jac72a, Jac72b, Jac73b, Jac73c, Jac74d, Joh95, JCKG^{+}50, Jon93a, Kat73, Kat92, Khr94, Khr96b, Kis93c, Knu89, Kom90, Kra78, Kra92, Kum95, La 79, LR79, Lan89g, Lan90m, Lan90n, Lap55f, LeR49b, Les93d, Les93e, Lew49, Lew63b, Lew71a, Lew72d, Lie46, Loc91b]. Report [Loc91a, Loc92a, Loc92b, Loc93c, Loc94, Loc95, Lon85b, Mab93, Mar89d, Mar98a, McG90, MF63, Men91b, Men91c, Mer46a, Mes70, Mey49a, MW91, Mil48a, MB48, Mil81, Moo96m, Moo97o, Moo97p, Mos93a, Mos93b, Mti96a, Mti96b, Mye92, Nel92c, Nel93b, Noy47, Obe98, Ore60b, Ore61b, Pan81, Par66a, Pat90a, Pea91, Pic51, Ple91, PI92, Pow47, Pra92, Pre92, Rau92, Rev92, Ros89b, Rot91b, Rot93w, Stu48, Sci51, Sci46, Sch68e, Sch08b, STFK09, Sch05c, Scl79, Shi46b, Shi54d, Smi55, Smi92d, Smi93d, Soa96, Spa96, Spi75, Ste91b, Str93, Sut95, Sut96b, Tam73a, Tar93, Tel46b, Tod79, Tur50, Tur51, Tur52, Tur53, Uni47, Ure46c, Veg97, Wee98, Weg96b, Wei52b, Wes98, Whi48b, Whi48a, Wol93, Woo96c, Woo97b, Woo97a, Woo97c, Woo97d, Woo97e, Woo97f, Woo98a, Woo98b]. Report [Woo98c, Woo98d, Woo98f, Woo98e, Wri97, Wur91b, Wur92, Wur94c, Wur96c, Wur96b, Wur97, You61, You63b, ZC92, Zam92d, Zam95c, Zam96b, Zhe90b, vH93, Ada86c, Ano49c, Ano82p, Ano88g, Ano91-34, Ano91-39, Ark85b, Ark85c, Ark85h, Ark85d, Ark85f, Ark85e, Ark85g, Ark85i, Ark86c, Ark86g, Ark86d, Ark86a, Ark86b, Ark86i, Ark86e, Ark86f, Ark86h, Ark87c, Ark87d, Ark87e, Ark87f, Ark87g, Ark87h, Ark88b, Ark89d, Bar77c, Bar78b, Bar80b, Bar81a, Bro80a, Cai89, CM12, Ekl77, FR76, Flo88, Flo89a, Flo89b, Isa85b, Isa85c, Isa85d, Isa85e, Isa85f, Isa86a, Isa86b, Isa86c, Isa86d, Isa86e, Isa86f, Isa87c, Isa87f, Isa87e, Isa87d, Isa88a, Isa88b, Isa88d, Isa88c, Isa88f, Isa88e, Isa89i, Isa89j, Isa89k, Isa89l, Isa89m, Kre86b, Kre86c, Kre86e, Kre86d, Kre87b, Kre87e, Kre87c, Kre87d, Kre87f, Kre88d, Kre88e]. report [Kre88f, Lau91a, LC78, Lon87, Mag88b, McC86b, Nat85c, NCA88b, Pai85f, Pai85c, Pai85d, Pai85b, Pat84, Ros88, SFT11, SF12, SF14, Sok22, Spe87b, Str56, WB87, Wal87b, WMRK79, vH77b, Ano53t, Ano57a, Ano60l, Ano70a, Ano85t, Bro57e, Lan80, Lap59g, Lap60a, Pat78c, Rab46a, Tel46c, Fra83, Lon85c, Ano80k, Rod70, Smi54c, Tho80b, Upt63]. Reported [Ano46-74]. Reporter [Fin55, Met83, Dro22g, Dro22h]. Reporters [Coh53b]. Reporting [Ano11d, Ano12d, Klo86, Lon86b, Kri95, Moo14]. Reports [Alb98b, Ano46-82, Ano47z, Ano48i, Ano48e, Ano49z, Ano50-36, Ano61o, Ano61p, Ano63u, Ano63r, Ano63s, Ano63t, Ano70d, Ano70e, Ano70f, Bro62a, Cal63, Cla66, Cof62a, Cof62b, Cof62c, Cof62d, Cor69, Dou68b, Dra73, Fed67b, Fel83d, Foz61d, Gla56b, Gol62b, Gre62, Gus70, Ham63, Hav69, Iff67, Jac67, Jac68, Lal64f, Lal66g, Lal66h, Lal66i, Lal67f, Lew65c, Lew71e, Med47, Mar64c, Mar64d, Mar64g, Mar64e, Mar64h, Mar64f, Mar64i, Mar65e, Mar65f, Mar65g, Mar65h, Mar66b, Mar65i, Men69, Mos89, Mur68, New62b, Oli66, Osb58, Pag68, Pre63, RH65, Rau69, RF65, RR68, Sub70, San91, Sch63a, SJ68, Smi71, Sot67, Sto66b, TT69, Tel66b, Tod64b, Tur60, Wil67b, Zei64, Zhe90c, Ano86e, Dew53, Rep83, Tho80g]. Repository [Ano71g, BM22, Ial18, Kla19]. Reppy [Dow81a]. Representation [Isa96b, Bla88, Fin90]. Representation-without-taxation [Isa96b]. Representatives [Ano45d, Rod61, BBG^{+}85, Gaf86]. repression [Kla79]. Reprocess [STC74, MvH97, Sai98]. Reprocessing [Coo06, DRM^{+}09, Jan15, RR15, Ram97, Zha15a]. Republic [Ano80j, Esp72a, Zho79, Smi67, Wes03]. Republican [Eav19, Isa87b, Mar64d, Rab48b]. Republicans [Ano60p, Cir00, Isa95e, Mec19a]. republics [Mot89]. Republished [Ano46-82]. reputation [Eps87a]. Request [Ano46-84, Pet78, Wes71]. Requests [Ano48r, Gar54, Mus65, Tho00]. Requiem [Lew71b, Mat82b, Moo05a]. required [Bar77b]. Requirements [Ano55w, Kna63, Ram64]. requires [Arn80, Arn22, Bec79, Cor78, GK21, Hun85c, Mec20c, Pan23]. Rereading [PF82]. rerun [Wie87]. Rescinded [Cal57, Dub57a]. Rescue [Fer68, Fly01a, Mar90, Pra14, Wil92c]. Research [And72, AL62, Ano46-68, Ano46-76, Ano46-73, Ano49-32, Ano50u, Ano62-28, Ano68h, Ano70g, Ano78n, Ano83q, Ano87q, Aur46, Aur72c, Bar77c, Bar78b, Bar79a, Bar80b, Bar81a, Bar48, Bea67, Ber72a, Ber66a, Ber49, Bla67, BFW47, Bru55b, Bus47, Cle48a, Coc51, CS63, Cre74, DZA47, Ded57, Ded59, Ded62b, Dew55, Doo64, DuB46a, DuB69b, Eis72, Eme60, Fed58, Fed67b, Fel83d, For80, Foz61h, Foz61i, Foz62f, Gar66, Gat72, Gol49, Gro53a, Gro53b, Has60, Hei47, Hux47, Jac71b, Jac73d, Jac73e, Jac73f, Joh65b, Jon70, Kay66, Kis66, Kno66, Kou75, Kow49, Lal65d, Lew70a, Lov63, Mar83b, Mas58, McE70, Mil63a, Nat56, New67, New58, Nor63, Nov77a, Opp47b, Osb58, Par73, Par62, Pic51, Pla60, Pla64b, Pow47, Pri69, Rab66a, Ric53, Ros68b, Rot79a, Sac48, Sac72a]. Research [Sch68b, Sea01, Shu78, Sim46a, Sin63c, Sin65, Sko67, Smi75a, Spi63, Ste46, Stu69b, Tee65, Tsi83d, Tuv53, Vol74, Wal51, Wat53, Wat57, Wei70b, Wen53, Wie47b, Wol55, Woo61, Wri63b, Zho79, dF79e, AM21, Ano77l, Ano87y, Ano91-45, Ano14f, Atk78, BDE^{+}77, Bal78, Day77g, Dro23a, Edw12, Gof93b, Gof93c, Har89a, Kap76, Lon86a, Mab93, Mak24, McC85, Sch84b, Sha87c, Sut84, Wol86, YG77a, Yud85, Ano80j, Dje68, WC47, Kat52a, Rot62, Wyl56, Ano89b]. Researchers [Sch05f]. Reserve [Ger49, Sin81]. Reserving [Ano90-44]. Reset [Ano72i]. Resettle [AO98]. reshapes [Isa91n]. Reshaping [Dou15, Myr66, Sha85c]. Residuals [Sue61]. Resigns [Ano49m, Ano49-31, Ano55f]. Resilience [Jor18, HK18, MD18]. resilient [Rot18]. Resistance [Ano68d, Pop19, Tau85]. Resistant [Act09]. Resisting [Win98]. Resnikoff [Wal84]. Resolution [Ano47a, Ano53w, Ano54h, Ano55t, Ano66s, Ano84d, BS57, Hum59b, Kih47f, Tho57, Kre88a, Saa84]. Resolutions [Ano46-83, Ano46-84, Jac62b, McC54a, Sip86]. Resolving [Bol63, Red82]. Resource [Ano84m, Ano86i, Ano89g, ACH84, Ber75c, CR71, Kon67, Scu65, Tod79, Wat57, Han96b, RW76, Wil77a, Har83, Dal82, Tod79, SD83]. Resourceful [EE85]. Resources [Alb73, Car49a, Eck78, Fit51, Gat72, Gor51b, H.84b, Kat51a, Kil62, Lal66b, McK74, Mei70, Mye87a, Rab51m, RE68, Sax51, Sin76, Smi73, Abr84, Ano76b, Coo78, Bal79b, Bre74, Gin54, Mei52]. Respect [Bra58]. Respond [BBZ01]. Responding [BBP14, Fel83d, Fis62b]. Responds [Fel69a, Sch68a, Ber87b, Gaf86, Tsi83f]. Response [Ano63-28, Ano70d, Bet77a, Car72, Cas77c, Com75h, Com72, Day75l, Day75m, DeN75, Dje73, EE85, Eme54, EB73b, Fry58, Gre73a, Ham93, Hol56, Kap78a, Kue61, Lev98, Lev06, Lov74a, Mar98b, Mul74a, Pal64b, Pan98b, Pri75a, Rab62i, Sim87a, Tam73b, TB70, Tra72a, Tra72b, TM98, WC74, Wyl56, ZSA73a, ZSA73b, ZSA73c, ZSA74, Al 15, Ani15, Ano15c, Ark87e, Bac12, BG14, Bin15, Bis14, Bor14, Cha16, DR14, Den82, Dha16, Esp14, Etz77, Eze16, Fis86b, FK12, GB11, Ger13, Ger20, Ger21, Gra16, Gub15, Gue14, Gya14, Haf84, Har14a, Har14b, Hue12, Hun14, Huq16, Ibr13, Jan15, Kar14, Kaz14, Ken80, Kro15, Lan14, Len14, Mal14, Mia15b, Mor14, Mor12, Nag15, Nie14, RR15, Riq15, Rof16, Sai14, San16, Sha14, Sha84e, Sin14, SW16, Thr15, Tom14, Tra90]. response [Ulg15, Wan16, Yin15, Zer13, Zha15a, Zha15c]. Responses [BH65, KKHK79, Hil12, Kan77a, Ram13a]. Responsibilities [Ano51-29, Ano53-37, Bro56b, DS49, Dub56b, Hil46, Hun87, Rab59j, Rus60, Tel47a, Tel63, Uni57, Wea63, Atk78, Fra47, LP85b, Pug78, Suz11]. Responsibility [All51, Ano88e, Ant50, Aug56b, Bri48a, Car76, Com57c, DS47a, Dar57, Das66, Daw68, Dub56b, Foz62g, Fra58, Gol48, Gor71, Gro56a, Gus85, Had56, HFO63, Hei70, Hoa72, Lon69b, Mas48b, Ols63, Pas55a, Rab48a, Rab48h, Rab56g, Rab69i, RD75, Rit48, Tru47, Ure48b, Ure48c, UM67, Wes55, Wie48, vL48b, Bal84c, Bey85, Gar85, Jon84, Kis76b, PFW18, Rog88b, Sob21, Wei76b, Ano88e, Gen84b, Ash48]. Responsible [Nus79, Oze92, Met24, Nuc20]. Responsive [Rei78]. Rest [Ing63b, NA92n, Nor02a, Ano93w, Car97, NA92h, Pan84]. Restiveness [Kug69]. Restivo [Lyt79]. Reston [Hog68]. Restoration [How65b]. Restore [Ano48n]. restraining [Mec21c]. Restraint [Fis61b, Rod69, Smi75d, Ste95, MFN17, Sha10, Tar23, Zha12]. Restraints [Rus70b, Sei69]. restrict [NK86b]. Restricted [Ano46c, Dro22a]. Restricting [Sto76a]. Restrictions [Tol58]. restrictive [Row77]. Restructure [Kok88a, Wei76a]. Restructuring [Le 77, DG07, Lyn88]. results [AK95b, Arn98a, Bar80c]. Resume [DRM^{+}09, Han98b, Hum60, AS85]. Resumes [Sch62c]. Resumption [Ano61p, Ken60a, Ste60c]. Resuscitate [Kib10]. Retaliation [Kin74, Sel63]. Retarded [Sch60b]. retells [Her90b]. retention [Ach12]. Rethink [Gro01, Zik21]. Rethinking [Brz82, Gar97a, Mey08, RAS04, Bus85, Ack85]. reticence [Spr25]. Retire [AGZ91]. retirement [Wei94d]. Retooling [BBC^{+}98]. Retracts [Ano48s]. retreads [NA88o]. Retreat [Boy89b]. Retreats [Ano69e]. Retroactive [Wei78b]. Retrofit [Mab93]. Retrospect [Woh54, Fel78h]. Retrospectives [She85f]. Rettig [Ano78k]. Return [Hit07, Man87, Mes71, Oph73, Arn92c, NA90l, Dow82f]. Returns [Ada06, Rab54g, Mil12a]. Reuben [Gre55a, Kee56, Ure56]. Reunification [Lal65b, Mul66]. Revamp [Gro01]. Revamped [FP90c]. Revanur [Dro23d]. Reveal [McM49b, BV20]. revealed [NA90j, NA93o, NA93s, Ano87-35]. Revel [Rab73d]. Revelation [Bou70, Cla59, Vis59, Day93, Mor59a]. Revelations [MR90]. revenge [Rot95f]. Revenue [Sta62b]. Reversal [New62b]. Reverse [DeN76a, Sch07a, Bro79a]. Reversed [Pau52a]. Reverses [Ano50q]. reverts [Isa93g]. Review [A.52, Abb84a, Abb85, Abb91, Abb93, Ach10, Ack85, Ack87a, Ack95, Ack97, Ada52, Ahe91, Ale69, Alf84, All55l, All55b, All55a, Alv87, Amr51, And72, And83, And55, Ano48y, Ano50-41, Ano51r, Ano54k, Ano77l, Ano81c, Ano87s, Ano05b, Bal79b, Bal88, Bam85, Bar79b, Bar79d, Bar80m, Bar88, Bar90b, Bar90c, Bei84, Bel93, Ber84a, BH62, Bet75b, BD79, Bir87, Bla82a, Bla85a, Bla87a, Ble87, Bou98b, Bou02, Bou87, Boy87, Boy89b, Boy91, Boy95, Boy05, Bra89, Bra87, Bre87, Bre93, Bri89, Bri80b, Bro55a, Bro55b, Bro79b, Bro47d, Bro81, Bru55a, Bur90c, Cah94, Cal82, Cam54, Cam89, Cap98, Cha68a, Che88, Cog82, Coh69, Col84b, Con81a, Con88, Coo81, Cop65, Cor94, CF46, Cra83a, Cra84d, Cra84c, Cra88, Cum98, Dal82, Dar91, Das96, Day75b]. Review [Day88, Day97, DeB81b, DeB81a, Der88, Deu83a, Dev69, Dha90, Dia83, Dia84, Diw79, Don86a, Don89, Don90, Dow86, Dow88, Dow79b, Dow79e, Dow79d, Dow79a, Dow79c, Dow80e, Dow80c, Dow80a, Dow80b, Dow80d, Dow81a, Dow81b, Dow81c, Dow81d, Dow82a, Dow82c, Dow82b, Dow84a, Dre82b, Dug91, Dun90, Dys72, Eas85, Ede89, Ehr79, Ehr82b, Ehr82a, Ehr83, EE85, Eld92, Eli85, Els90, Eng91, Eng88, Eps75b, Eps90c, Eri95, Eva86, Fel81a, Fel83a, Fel86a, Fel92, Fel83e, Fin90, Fis55, Fis90, Fla90, Fla80b, Fly98a, For83, For87, For86, Fra81, Fra82, Fra91, Fre84b, Fre84a, Fre88, Fre95, G.84b, Gai69, Gai92, Gal86, Gan82, Gan83b, Gan83c, Gar83a, Gar54, Gar97a, Gen84b, Gen84c, Gil83, Gil85, Gle94, Gol92a, Gol93a, Gol88a, Gol88b, Gol95a, Gol96a]. Review [Gol96b, Gol93d, Gol95d, Gol83c, Gol88g, Gom84a, Goo84, Got87, Gov93, Gra93, Gre87, Gro01, Gue80, Gus85, H.84b, Hac81, HD82, Haf79, Hah61, Ham61c, Ham91c, Ham93, Ham83, Han90c, Han92b, Han94, Han95a, Han98a, Har85b, Har79b, Har81, Har86e, Har92b, Har92a, Haw50, Her85, Her88b, Her90a, Her90b, Her82b, Hir88, Hob85, Hoe88, Hol83a, Hol86b, Hol86a, Hul98, Hun85a, Hym91, Ing82, Ing79a, Iri79, Iri86, Ise92, Jac73d, Jac73e, Jac73f, Jac80b, Jac92, Jer87, Joh81a, Joh97a, Joh00a, Joh10, Jon63a, Jon87, Jon90b, Jor87, Kal68, Kap91a, Kar80, Kat52a, Kat52b, Kee56, Kel98a, Keu83, Key81, KFM50, Kin80c, Kin80b, Kli81, Kli90, Kor92, Kra94, Kre83a, Kre85a, Kre90, Kre93, Kre89c, Kre98, Kus83, LaF90a, LaF86, La 91, LL90, Lad81, Lan55]. Review [Lan98, Lap55a, Lap61a, Law88, LeR50a, Leb91a, Leb92, Leb94, Lei92, Lep89, Ler80a, Lev84b, Lin84a, Lin85, LP92, Lof51, Lon82, Lon85c, Low93, LD79, M.84, Mac84b, Mag81, Mar69, Mar91a, Mar61d, Mar85a, Mar89e, Mar94, Mar97a, Mar98c, May89, McC93, McC96, McE88, Mea93, Mei52, Mei90, Men92b, Men91d, Mer79, Mer90, Mes82, Mes83b, Mes84b, Mes84a, Mes85a, Mes87, Mil85a, Mil88, Mil85c, Mil85b, Mil91, Moo93a, Moo93b, Moo94a, Moo95d, Moo95b, Moo95c, Moo95e, Moo96b, Moo96c, Moo97a, Moo97b, Moo98b, Moo00a, Moo02d, Mor82a, Mor79, Mor83b, Mor84a, Mor87c, Mos92, Mos81b, Mos81a, Mos90, Mos91, Mül90a, Mur84, Mye88, Mye89, Mye93b, Nad98a, Nas82, Nat86b, Nat87, Nor88, Nor02a, Nor04, Nor06, O’C87, Oze92, Pal02, Pap88, Par73, Par80, Par46b, Par88b, Par86, Pat79a]. Review [Pat82a, Pat83, Pei80, Per91a, Per94, Per95, Per87b, Pit91, Pol93, Pol91b, Pow88, Pow89b, Pow90a, Pow91b, Pre86, Pre96, PvH72, Qua80, Rab48c, Rab53f, Rab98, Rad83, Ram88b, Ran85, Rap11, Rau91, Ray81, Rec89, Rep85a, RK87a, RK93, Rob91b, Roc85, Rom82, Rom89, Rom90, Ros90a, Rot89a, Rot93a, Rot96b, Rot97b, Rot97c, Rot98a, Row88, Sac84, Sag86, San90, Sar85, Saw79, Saw80, Saw81b, Saw81a, Sca91, Sch90, Sch81a, Sch86, Sch69d, Sch80e, Scl79, Sco79, Sei81, Ses83, Sha90, Sha50, Sha84b, Sha91a, Sha91b, Sha81a, Sha95b, Sha98b, She82, She83b, She83a, She89, Shi54c, Shi84, Sho92b, Sie85, Sil79a, Sim84b, Sim87a, Sim85, Sim88, Sim56, Sim90b, Sim11, Sin81, Siv80, Sla88, Smi90a, Smi85d, Smi88, Smi97, Sol80a, Som84b, Som91, Som84c]. Review [Soy89, Sta85, SB91, Ste91a, Ste89e, Ste80a, Ste86c, Ste95, Sto55a, Sto55c, Sto55b, Sto85, Stu89, Swe88, Swe89a, The89e, Tay87, Tay91, Tay98, Tel49, Tho79a, Tho80c, Tho80b, Tho90, Tic94, Tir88, Tod80, Tra90, Tsi89a, Tuc97, Tum87, Uhl95a, Uhl95b, Uhl96, Uhl97a, Uhl97b, Uhl98a, Uhl98b, Ure56, Val86, Van87, Van91, Van90b, Van03, Van88, Ven80, Wal81a, Wal85a, Wal87a, Wal83b, Wal84, War86, War92a, Wea82, Wei85b, Wei89a, Wei89b, Wei48, Wei55b, Wel79, Whi89, Wig62b, Wil81a, Wil92a, Wil80a, Wil80b, Win84b, Win90, Win94, Win98, Win81a, Woh54, Woo92, Yar80a, You82, Yud83, Zag91a, Zel92, Zim81b, vL92, Ano10a, Bar80c, Bar80k, Eps81, Eps90b, Fin77a, Fin77b, Flo89a, FK12, GG11a, Kib12, Pug80, SS90a, Rot79a]. Reviewer [HY67]. Reviewing [Pab88, HP80]. Reviews [Ada87, And78a, Ano62k, Ano78e, Ano78d, Ano78f, Ano79a, Ano81f, Ano98h, AM86b, Ash81, Asp79, Aue78, Bal79c, Bas80, Ber79, Bla84, Bri86, Bus85, Cal84a, Cha79a, Chi78, Cra83b, Cra84e, Dav79b, Deu83b, Div79, Don86b, Dow80g, Dow80f, Dow80h, Dow80j, Dow80i, Dow81g, Dow81f, Dow81i, Dow81e, Dow81h, Dow82d, Dow82f, Dow82e, Dow82g, Dow82h, Dow83b, Dow83c, Dow83e, Dow83a, Dow83f, Dow83d, Dow84d, Dow84b, Dow84c, Ehr81, Ekl80, Fel83d, Fis83, For80, Fra83, Gen86, Goo86, Gre82b, Gro78a, Gus83, Hal93a, Har88c, Her83a, Hol91, Hos87a, Hun87, Ing79b, Jas81, Kel79, Kil78, Kin86, Klo85, Kri85, Lac85, Leb91b, Lyn78, Lyt79, Mad79, Mar79a, Mar79b, McC88a, Mes83a, Met83, Mil86b, vCM99b, Mor85a, Mye87a, Nad79, Nei82a, PMRP50, Pat81a, Pos79]. Reviews [Rea80, Rep86, RK89a, Roc81, Roc83a, Rod85, Rot79a, Say84, Sch80c, Sch79b, Scl79, She79, Sig85b, Smi83, Sti83, Swe86, Swe93, Tho80d, Tod78a, Tod79, Tod81, Whe84b, Wil71, Wil84, Woo82, Wou81, Wun82, Rob20]. Revised [Ano46-85, Lev47, Lev48, Gar88b, Mey94, Mol62a, NA87i, Sha54b, Tra71]. revision [Flo89b]. Revisionism [Mey94]. Revisionists [Pra03c]. revisit [Kor21]. Revisited [Coo62a, Gre73c, HE72, McN63b, Rab72d, Ack90j, Ano87-27, Dre84a, Fel77n, Isa86b, MW83, Osg85, Sha78a, Sha78b, Yos80, Phi84a, Roc81, Mad79]. Revisiting [GM10, Spr19]. Revitalizing [Tur13]. Revival [BM08, Kem09, Har87e]. Revive [Alb87a]. Revived [WWHV07]. revives [Lea87]. Reviving [Ano98o]. revoke [Gol21]. Revoked [Ano61m, Pod24]. Revolt [Eps96b, Law69, Rab69k, Bla97]. Revolution [Ano56f, Apt61, Bin79, Bor71, DeB81a, Gol69b, Lib01, Nix60b, Phi54, Pie59, Rab63g, Rab63h, Rab63f, Rab63i, Rab68e, Rab73d, She61, Ste91a, Tra71, Wei56a, Wyl69, Ano69d, Ell77, Gun20, Ham91b, Kre96, Lif13, Mat18, Mec19b, Mec20f, Mec21b, VU22, Wei78g, Ano69c, Ano69s, Cha69a, Git69, Jac74b, Lyn69, Rih65, Sar72, Kap91a, S.80, Moo96b]. Revolutionary [Cic74, Ehr81, Fit69]. Revolutionists [Mil70]. revolutionizing [HD23]. revolutions [Pea92c]. Rewards [Cha58a]. Rexford [You56]. Reykjavik [Alb87a, WH11, Rho09]. RFK [Fie24]. Rhetoric [Men91a, You75, Cas78b, Kin81b]. Rhine [Mul64b]. Rhoda [Sut54]. Rhode [Rho09]. Rhodes [Han95a, Mes87, Moo95d]. Rhyme [SG65]. Riaz [Men92b]. Rice [Oga58, NA91g]. Rich [Lef76, Nas62, Sal63a, Rab72a]. Richard [Abr73, Ada87, All54d, Ano78d, Ano78k, Ano78n, Ano86k, Ano90p, Bla84, Bla85a, Ble87, Bre73, Dow80g, Dre82b, Eml66, Eng88, Fol75, Gol69b, Gre71, Han94, Han95a, Hol83a, Hun85a, Ing70, Jac80b, Keu83, Kra94, Lal67a, Lam69, Mar78a, Mar79b, Mes87, Moo95d, Mul65, Nad98a, Nie63, Pla57b, Rob91b, Ros90a, Sha91b, Sha98b, Sig85b, Wil84, You82, Zel92, dF79b, dF79a, dF79d, Lug00, Rho09, Bul09i]. Richelson [Mos90, Row88]. riches [Isa94i]. Richland [McE52]. Richter [Ano12b, Dow79a]. Rick [Dow83d]. Rickettsiae [Ang60]. Rickover [Lap54c]. rid [LC94]. Riddley [Mes82]. ride [Bro20]. Ridenour [Ada47, Bus47, Ein47, Ell47, Har47, Hux47, Mer47, Mor47b, WC47, Wie47b, You47]. Rides [Mar64i, Cha78b]. Ridge [Ano46-69, Ric74, Ano48t, Ano49-33, Doy17, McE52, Nor47, Whi48b]. Ridgeway [Mil69a]. ridiculous [Bel20a]. Riding [Sch04b]. Riehlman [Shi54d]. Rieser [Moo99g]. Riga [Fel77o]. rigged [AO95b]. Right [Chr71, Dud75a, Hol05b, Isa95e, Isa95f, Len76, LH99, MM77, Moo03a, Opp90b, Sch74c, Bul09e, You69, Al 15, AA54, Ano94n, Ano95n, Ano98q, Bin15, Bli76, CP98, Dar91, Dro22f, HBD17, Isa97f, Mec19d, Mec20d, Moo92g, Pie22, Rot97f, Sol80c, Spa95f, Tsi82b, Ulg15, Isa96d]. Right-wing [Isa95e]. Rightful [MT03]. Rights [Ano53-37, Atk78, Dro22a, Fel83d, GL48, GA08, Hel77a, Lis46b, Mac83, Smi69a, WG69, Yoa78, Zim79, dRMT^{+}47, Aja78, Ano88u, Ber82, EV84, Eva82, Fri84, Fri79, Gol82a, Gra86, Hai86, HFMS79, Kan81, Lop76, Mel78, Pri81b, Sco92, Smi84a, Ste81b, Ste82b, Von83, Zim78, vL86, Bok49, Gol78c, Woo82, Ano78p]. rigid [Ano91-49]. Rigor [End00]. riles [Spe87b]. Rings [Dow84d]. Rip [WF16]. Ripe [Flo89b]. Rise [Cow65, Cre70, Dow81f, Eps75c, Gol92a, Hec70, Jer87, Kin69, Mar70, Mar71b, Mil69b, Rab72a, SM89b, Van87, Wil73a, Wit01d, dJ63, Boy95, BR15, Dro24c, Isa87f, Jac15, McL18, Moo92a, Nas18, Nuc18, Ros81, Sti18, Tre18, Van20]. rises [Pai83c, PP83]. Rising [Ima93a, Kor17, Lay07, Fie18, Gut93e]. Risk [And78a, Ano54b, Ano82f, Ano85p, Coh53a, Dav48, DWH83, Dow82a, DuB55b, Fin90, Gol74, Knu89, Lan89g, Mat57, Mav75, Rab54f, Rab54i, Rat69, Sez16, Sim84b, Szi54, Wea75, Wil81a, Akh17, Ano84n, AH22, Ban11, Ble88, Bor87, Car17, Cle92, Con90b, DM18, Eav17c, Ehr82c, Fer10, Hil12, Hol83a, Jac15, Kas12, Lym11, Mes97, Pan22, Pod16, SMĆ08, Sar17, Slo12, Suz11, Ano78k, Hol83a, Hol83a, Scl79]. Risk/Benefit [Hol83a]. Risks [Ano63r, Cha58a, DeN74a, GM09b, Gou61, Joi54, Lac57, Lin85, Mac77c, SD10a, STC74, Ste72a, Tho80g, Ahe90, Ano11c, Bey12b, DS47b, DM24a, DM24c, Fie18, Flo78, HV18, Lef77a, McL18, Mil12b, PFW18, Rot78b, Sah20, SD17, Tor19, Tra24, WH16, Wel80]. Risky [Ano91-39, Per87b, Wes01, Hel11]. Risse [Ano89g]. Risse-Kappen [Ano89g]. Ritchie [Deu83b, Fer57a, Wag61]. Rites [Dia63, Moo97a, Ste89c]. Rittenhouse [Mer79]. Ritualized [Mic61]. rituals [Hen84]. Rival [Voh71c]. rivalries [GN10, Ano10a]. rivalry [Cou77, TS22]. Rivals [Soa96, Kli90]. River [Rot89a, Ter64, VSD^{+}09, Gor92, Zag94b, Lan90o, McL53, Mon53b, Pea99]. Riveter [Dow81e]. Road [Ano98h, Bos06a, Fry56a, Hig46, Kre87d, Lan96, Mun65, Old69, Sah20, Szi62a, Van90b, Zai06, BV10, Gol90e, Isa92i, NA89a, Pai85c, Squ21a, Van03, Van86]. Roadblocks [Ano59i, Rot92l]. Roads [Ran55b, Gan76]. Roald [Bou87, Moo94a]. Roaman [Ano78n]. Roared [Pai81a, Shu81, Swe81]. Rob [Wil81a, Dro24c]. Robbins [Ano78d]. Robert [Ano78n, Ano85b, Ano86i, Ano14a, Bet58, Bet67, BS05, Bru59, Day88, DeB81a, Dev69, Dia83, Dow86, Dow79a, Dow84d, Ehr82a, Eli85, Fel79o, Fer62, Fis83, For80, GNMS54, Gol55, Got87, GDH^{+}04, H.85, Han94, Har85b, Ing79b, Ing79a, Kal54, Kal63b, Keu83, Kin86, Liv69, McM05, Mea62, Mea93, Mes64b, Moo95e, Nic54b, Pow88, RB67, Rei78, Rob62, Roc81, She89, Sin81, Smi83, Sti83, Swe86, Tho80g, Uni54, Van91, Wea82, WS79, Wel79, Whe84b, Wig62b, Woo82, You65a, dF79f, dF79c, Ano65f, Ano85h, Mec22a, SW80a, SW80b, Wil82b, Wol91b]. Roberts [Jac73g, Smi85d]. Robin [Sid69]. Robinson [Tho80d, Bar90c, Tho80d]. Robles [Bar67b, Fel82g]. Robock [Ano15a]. Robot [Pra14]. robots [Gub14]. Rochester [Ano72a, Mar70]. Rock [Cop65, Ano95m, Lev12]. Rockefeller [Lap60b]. Rocket [Coo62b, Kel61, Mas58, Ros90a, Szi60a, She97]. Rockets [Ano56d]. rocks [Ano89k, Cha89a]. Rocky [Ack01, Ano46-90, Kre87d, OW15, Aba89, Ack94, AZL90b, AGZ91, NA89a]. Roderick [Bla84]. Rodney [DM24a]. Roe [Say84, Sch54a]. Roentgens [Ano55u]. Roff [Sca91]. Rogat [Swe88]. Roger [LD79]. Rogers [Ano78n]. Rogue [AH03c, Moo95b]. Rogues [Gab01]. Roiling [Lam87]. Rokkasho [Kat16]. Roland [Bou61]. Role [Ano74e, Aus60, Ber66b, Coo76c, Dea51, Dob54, DuB46a, DuB66, DB67, Eis68, Fel70a, Fel83d, For57, Gol75, GW65, Kin80b, Kra46, Lal65c, Mas48b, Mul65, Nel74b, Par66c, Per95, Rab57h, Rab71d, Sen57, Sai72, Smy50b, Spi67, Spo64, Ste74, Tel66a, Thi47, Ure48c, Wat51, Wei53a, Wei52a, Wie48, Wie62, Wri63a, Yut69b, vL48b, Abt84, AM21, Ano15c, AE85, Ark86a, FTvHW76b, Han17, Kul17, Lar50, Mcc87, Rab60e, Raj17, RH17, Sam22, Tsi77a, Van90a, YG77a]. Roles [Uhl98b, Bat86]. Roll [Ano92y]. Rolled [Coo71]. Rollins [Ano46-86]. Roman [Eva86, Dav57b]. Romance [Ben59, Pat90a]. Romania [Woo97f, Woo98f]. Romanovsky [Dro25c]. Rome [Cal61a, Cha73b, Han75]. Roméo [Sch06c]. Romm [Ano15e, Moo93a]. Ron [Dow83f]. Ronald [Lak66, Rab69b, Siv76, Smi83, Ano11f, Kre86d, LD89, NA88n]. Ronder [Dow80h]. Ronneby [Int67]. roof [Ano95n]. Room [Jac71a, Mye82b, Sim59e, Cav13]. Roosevelt [Hol06]. Roost [Sie05c, Isa92h]. Root [Rus51b]. rooted [Mcc88c]. Roots [DeM76, Fil76, Mau77, Pan71, Wun82, Fel77o, Kas89, Lak78, Sch19]. Roper [Dow84b]. Rose [Kri85, Mes85a, Sha99, Sig85b, Ano11g]. Rosebury [Has71, McL49]. Rosemary [Cal82]. Rosen [Sti83, Sto56a]. Rosenberg [dF79h, Mar53c, Mon53a]. Rosenbergs [Gla75a, Sob91b]. Rosenblum [Ano85d]. Rosenfeld [All56l, Hen71a]. Rosie [Dow81e]. Rosner [Dro24b]. Ross [Sha81a]. Rostow [Dia84, Isa81a]. Rotblat [Ano98h, BM99, Moo93b, Smi58f, Ano92-32, Day96, Fei77, Lan96, Soc54]. Rothman [Rom90]. Rothmann [Bru50]. Rothschild [Mes64a]. rotor [AH91j]. Rotten [Gro07]. Rough [Hol05a]. Roulette [Mes83b, Rot92q]. ’round [Alb98c, Ano63t, Bro88b, Eve88b, Gro52, Isa91k, Lap59e, Ram64, Ros75, Sto67, Day75a, Dea89, Fel80h, PE89, vH91]. Roundup [All54c, All55c, All55d, All55e, All55f, All55g, All55h, All55i, All55j, All55k, All56a, All56b, All56c, All56d, All56e, All56f, All56g, All56h, All56i, All56j, All57a, All57b, All57c, All57d, All57e, All57f, All57g, All57h, All57i, All57j, All58b, All58c, All58d, All58e, All58f, All58g, All58h, All58i, All58j, All58k, All59b, All59c, All59d, All59e, All59f, All59g, All59h, All59i, All59j, All59k, All60a, All60b, All60c, All60d, All60e, All60f, Ano54s, Ano54t, Ano54u, Ano54v, Mar60a, Mar60b, Mar60c, Sim54e, Wid60, Wid61c, Wid61d, Wid61e, Wid61f, Wid61g, Pai82g]. roused [Sim87b]. route [NF13]. Routine [Ack88g]. Row [Wil71, NA91g]. Rowan [Mos81b]. Rowe [Ano78k, Bul08m]. Rowny [Isa81a]. Roy [Ale69, Ano78o, Ano86l]. Royal [Bro20, NA93m]. RRW [Gus07]. rubber [Sol80b]. rubbish [Abb84b, Pen84]. Rubble [Rot01c, Wur91b]. Rubin [Gri83a, War86]. Rubinstein [Ano86i]. rubles [Hol80b]. Rudolf [Fel86a]. Rufus [Ano78n]. Rugged [Dra69]. ruin [McM05]. Rule [Ano46-46, Ano68j, Dam62, Ken77, Kre03, Ste91c]. Rules [Kre05, Zha99a, CM91, Fai80, Gof93b, Gof93c, Joh94e, Nol90, Row77]. Ruling [Lon85c, Shi47a, Wei86c]. rumble [Rot95m]. Rumbles [Isa98c, Wie67]. Rumsfeld [Gar98, Isa98c]. Run [CB95, Coh02, Gar00a, Rot93q, Dow82h, Kre88c, NA88g, Rot93e, Str94, Zim87, Fox48]. Runaround [Ano90-32]. Runaway [Tyn75]. Running [Ano91-36, O’B69b, Pai81b, Sch92a, AH91f, Ano97m, Man21, Wou81]. runs [Ano92w]. rural [Mak76]. ruse [Win87]. Rush [Cir98, Day75h, You66c, Rot97k, Jac73c]. Rusk [Sha91b]. Russel [Whi89]. Russell [Ano85b, Fer57a, Ing56c, Kli81, Mes82, Sto68a, Ano80o, Kap61, LB68, Mar58c, Rot81b]. Russett [HG71, Rus71]. Russia [Kus83, Mar79b, Rab71a, Uhl96, Ano46-87, Ano94v, Ano96n, Arn54a, AK01, BC16, Bro49, Brz12, Bud22, BBC^{+}98, BHK^{+}21, Cav92, Che52, CJ01, Chu04, Chu50, Jr22b, Dro22h, Dun48, Dvo07, Eli46, Eva99, EO00, Fly01c, Fly02c, Foz61i, Gei16b, Gol93c, Gro49, Han96a, Har22, Hol97, Hut46a, Iln97, JCF02, Khr94, KS98, KK02, Klu22, Loc93a, Lor01d, Mec16c, Men95, Mye93a, NA93f, Osb49a, Pik06, Pod01, Pod24, Pop93, Pos00, Rab52i, RK19, RMS03, Rot97h, Sei49, Smi94b, Sti48, Sut16, Ure49a, Vad21, VS04a, Wes03, Wol01a, Wu22, Yaf06, Zag95e, Zys17, Rab48c]. Russian [Ano14a, Cra83b, Mar94, McC88a, AH20, Ani15, APR17, Ano46b, Ano49-34, Ano50-42, Ano53-38, Ano93v, Ano99-27, Arb18, Arb19, Ark00d, Ben01a, Bet60, BV10, Bug22, CS16, Chi17, Chu94a, CNB95, Cra82a, Dob52, Dow92, Dro22g, Ekl80, Gan52, Ger58, Git74, Gro46, Gro47b, Han93, Isa93d, Joh94d, Kap94, Khr96b, Kli52, KK19d, KN11d, KN12c, KN13c, KN14c, KN15d, KN16c, KN17c, KN18c, KK20c, KK21d, KK22d, KKR23b, KKJK24c, Lon52c, Mab93, Mia15b, Mor12, NA93d, NA93n, NA93o, Nor94a, NA94a, NA94d, NA95b, NA95d, NA96d, NA96b, NA97a, NA98d, NA99c, NA00e, NAKH01d, NAKH02d, NK03c, NK04b, NK05e, NK06b, NK07d, NK08d, NK09c, NK09d, NK10d, OS00, Pan23, PB^{+}01, Ric52, Rot92q, Rot98b, Sag93, Sal16, Smi72a, Bul08c, Tel58, TPPB21]. Russian [Tre19, TM97, Tro20, Wil72a, You77a, Zag95c, vH93, Mes83b, Nor02a]. Russian/Soviet [NA93o]. Russians [Cra84e, Dow84d, Ano92x, Ano96p, Ash81, Cra84g, Dro22g, Got65, Lee91b, Llu00a, Llu00b, Mor50c, Osb47, She69, TM94a, Wor58, Dow81d, Kre93]. Russo [Szi60a]. Rust [Alb94e]. Rustamova [Dro22h]. Rustum [Ano86l]. Ruth [Ano90p, Ded62a, Fer53, Gol96b, Goo86, Kal68, Ano84o, Ano87-36, Ano87-37, Ano88r, Ano89x, Fel84g, Rep83]. Ruthless [You91]. Rutskoi [Gut93a]. Rx [Fin77a]. Ryall [Dow80d]. Ryukichi [Mos81b].

S [Abr73, Ano45n, Ano56g, Ano78d, Ano78l, Ano81c, Ano91s, Ano13c, Ano14a, Bea47a, Ber77b, Bro57b, Bru50, Cal82, Cav57, Cra83b, Cra84c, Don86b, Dow82f, Eml66, Fel92, Fol75, Gan83b, Gol95d, Got87, Gre58, Gre82b, H.84a, Han94, Han98a, Hob85, Jer87, Joh51, Kli90, Lam69, Leb91b, Lof51, Lon82, M.85, Med47, Mar58a, Mil85c, Mor85a, Mor87c, Mos81b, Mul66, Nas82, Nat87, Nor64, Olu71, Pan63, Par80, Rab67a, Ras55, Rez07c, Roc81, Roc83a, Saw80, Sha91b, Sim84b, Smi59a, Smi75c, Sno68b, Ste91a, Stu89, Tho80d, Tho80g, Tsi72a, Uhl95b, Wal85a, Wil76b, Win94, Win98, Wri63a, dF79h, dF79c, dF79g]. S.C.O. [Mat84]. S.I.O.P. [Lev84b]. Saaty [Dav69]. saber [Wei78b]. Sabine [Gol92a]. Sabotage [Flo76a, Lon85b, Sch74b]. SAC [Bai92]. Sachs [H.85]. Sacred [Cof67, Coh93a, Wil75a, Ano95l, Fei81b]. Sacrifice [Ber46, Mey08]. Sacrificed [Shi55]. sad [Bal89b]. Saddam [Ark00e, Ham98, Isa90y]. SAFE [Ano82l, Ano90-45, Bet77a, DS47b, Fel75c, Fly02d, Gul75, Jac74b, Kra99, Ste86b, dL59, Lug00, Ark98f, Bar87, Bro86d, Cha89c, Jag98, NA92e, Pan91, RS16, Sch11, TPPB21, Wie82a, DM82, Gue79, Riv96, Dow80h, Dow82h, Phe63]. Safeguard [Rat69]. Safeguarded [Ore63]. Safeguarding [Sch74a, Wil75b, Don86b]. Safeguards [FB89, Lan83b, Spi74, Wei17, You47, Bol88a, Bol88b, Bol88c, Ekl81, FM82, Fis82, Fis86a, Mog81, Pai85f, Ric81, Rot77, Swe78b, Wei78f, Ste76a, dKL83, DeN74a]. safely [Ano96m]. Safer [Lym08, AT92, BP90]. Safety [Alv87, Ano82f, Ano92r, Aue02, Bet75a, BGR61, Car94d, Far71, Hav04, Isa91l, Kep76, Kli76, Kou75, May17, McN09, NA88o, NA91h, PvH74, Pri75b, Pri75c, Ras75, Rau69, RF65, Rot02e, Sch75b, Tur53, Yon68, You75, vH76a, Ano77l, Bla87b, Car94b, Coo12, EG24, Gil18, GT14, Hig77, Kid91, KK12, NGG11, NA91o, Pan91, Pai82d, PI92, Sch11, Utr87, Zam92c, vH76b, vH91, Bet75b, Rot79a]. Saga [Isa04f, Sol80b]. Sagan [Rob91b, Smi85d, Ber75a]. Sagdeev [Moo94a]. Sagdeyev [Bou87]. Sage [Bro05, Nor06]. Sagittarius [Tel51b]. Saharan [Yud85]. Said [Cal60a, Gan72, Rab50c, Ano93v, Ano97l, CE88, DvH91]. Saigon [Day75e]. Sake [App53, Gre05, Rab69c, Wyl53]. Sakharov [Hol90, Hol91, Ano82s, Ano84e, BP85, Bet90, Blu84b, Cle71, Eva87b, Gor85, Hol91, Jan78, Jan80, Mar87b, Rab68c, Sak82, Yor81, vL85]. Sakharovs [Roe91]. Salas [Fly00a]. Sale [Mur68, Deu84a, NA93f, Pat82b, Rot93b, Sco90, Rot96b]. sale-ski [NA93f]. Sales [Ahm03a, Ano46r, Ano67b, Ben01a, Kum04, Wit01b, Ano86q, Ano87-30, BL79, Fer10, Fly97, Gor95a, Isa91k, Kla76, Mac91, Gan82]. Salinas [Gor96c]. Salisbury [Rab71a]. Salk [Ehr82b, Ehr82b]. Sallie [Cal82]. Salmon [Opp03b]. SALT [Dow81g, dF79a, Her78a, Lew71a, Sku72, HC71, Ano77m, Asp78, Bec79, BB79, Cal80, Cal73, Cof73, De 70a, Dow82i, Dre77, Ear86, Fel72b, Fel77i, Fel78m, Fel78g, Fel79j, Fel79g, Fla78, Gue79, Hoo81, Ind81, Ima71, Isa78, Jac79, Jac82c, Kat80b, Kin81b, Lap72, Lew70d, Man79, Mat79, Pug71, Per71, Pie80, Rat72, Sch81b, Sha80, Spi71b, Sto78a, Tro79, WN79, Ber73b]. SALT-II [Fel72b]. Salty [Sch79c]. Salvador [Bla82a]. Salvage [Mar71b]. Salvaging [Sie04c]. Sam [Mor83b, Wil81a, NA89f, Ano91v, Com74, Jr22a, Isa90y, Isa91q, Mas67]. sama [Joh96a]. Same [Ple91]. Samir [Lal67b]. Samizdat [Ano78o]. sampler [Car94d]. Sampling [Zei73]. Sampson [Kal78]. Samra [Moo95d]. Samson [War92a]. Samuel [Ano78f, Ano89f, Cap98, Eng91, dK60, Lap57, Mol62a, Mor62d, dF79c, Wei66a]. samurai [Dow91]. San [Fis81, Gei81, Mac81b, Mar64d, Sch74b]. sanctification [Wei85a]. Sanctifying [NK86a]. Sanctioning [Har22]. Sanctions [Ahm03b, CP93b, CA98b, Dav93, Dum93, Ela93, Ell93, Gra98, HM98, KOE93, LC93, Shi47b, Wer93, Woo93, Ano86f, HS13, Klu22, Wat19]. sand [Ano92w]. Sanders [Pre62]. Sandia [AC02, Ano49e]. Sandino [Bar67a]. Sandra [Roc85]. sandwich [Ano93z]. Sane [Sib73, Cor84, Fer88, KFT61]. SANE/Freeze [Fer88]. Sanford [Lew62a, Uhl97a]. Sanity [Ano85d, Ant69, CF46]. Sanskrit [Van60]. Santa [Pat90b]. Santillana [Sha54b]. Santoni [Dow83f]. Sapiens [Bor52a]. Sarabhai [Sea72a]. Sarajevo [Diz94, Moo94j]. Sarma [Dow79b]. Satellite [Ano55h, ER90, Mas58, Ric03c, Ric05, Smi71, CET77, JPR90, Pal77a, PPE22, Por19, Rob88]. Satellites [Ber58, Dav57c, Pie61, Pri58b, Rab57b, Rab67b, Sch67a, Smy61, War59, WJ61, KM89, Kre89b, Lin88, Ric89, Spe89a, Tir86]. Saturation [Szi64, Szi65]. Saturn [vB69]. Saturn/Apollo [vB69]. Sauce [Rab67i]. saucer [Rot98a]. Saudi [Sez20, Dro16b, Sez16]. Saunders [McK74, Mye87a]. Savage [Kli59]. Savannah [Hor88a, Hor88b, Lan90o, The88]. Save [Gol89, Rad92a, VSD^{+}09, Whi78b, vW58, Eav17a, Gol21, Gou24, HBD17, Hus24, Pos21, Sag93, Sch24, Spa95f]. Saved [Ber86b, Kah59, SK03, Her84, Hol94b, Kre88b, Mec20b]. Saving [Arb16, Fra90b, Ano91-41]. Savor [Cof73]. Saw [Sha91b]. Say [Ano93y, Bra58, Kri91, Sch59b, Sul58b, Ano96m, Cha84, Coo76a, DvH91, Gam94, Isa92n, Lan79, Mac92, Mar89c, Meh21, Ola82, Sha84d, SOE21, Vog90a]. Sayer [H.84b]. Saying [Ano91-56, Dys75, Ehr75, KOE93, Luc90a, Bas80, ET20]. Says [Ahm04, Jac74b, Joh01, Kib03, Mae01, Mei01, Orm57, Rab50c, She00, Zim89b, Ano88g, Isa95e, Isa95d, Men96b, Moo96n, Nel93b, NA89s, NA91m, Pat90b, Riv96, Rot92n]. Scale [Kow49, Moo96g]. scales [Rev21]. scaling [LBCA23]. Scandal [Sch99]. scandals [Alb89b, Brz89b]. Scandinavia [Lau91b]. Scandinavian [Bar80j]. Scant [Day74d]. Scanzano [Gun04]. scapegoat [Gol91a]. Scarcity [Ber75c, Mye89, Pal64b, AH22]. scare [Alb94b, Boy86a, Jon13, Lan90j, Rub90, Uhl98b]. scare-nario [Alb94b]. Scared [Gou74, SV86, Sau74, Sch05d, CW23]. Scary [Kos04a]. SCC [GK85]. Scenario [Cal73, Mul71, Net70, Sch74b, LJ17, vHW76b]. Scenarios [How65a]. Scene [Hin69a, Hin69b, Lor01c, Pri69, Spi58, Wei83a, Gru70b]. Schaeffer [Van91]. Schaerf [Sch79b]. Schapiro [Rab64a]. Scharre [Mec23c]. Schechter [Leb91b]. schedule [Lew72b]. Scheer [Dia83]. Scheinman [Hoe88]. Schell [Ano98h]. Schelling [Ber77b, BGR61, She61, Bul67, Cro68, Sch68a]. Scherer [Ull63]. Schild. [Ber79]. Schiller [SS67]. Schilling [Sco74, Wol63]. Schilpp [For80]. Schizophrenia [Bra02]. Schlesinger [Lew72c]. Schlossbert [Ano78n]. Schlosser [Ano14d, Dro23c]. Schmidt [Pos79, Bar88, Sel63]. Schneider [Hob85, Smi90a, Bul09c]. Scholar [Ano85a, Pol66]. Scholarly [Ros68b, TV71, Ano78p]. Scholars [Fei60, Fei61, Hes78]. Scholarship [Ken81, Ano82s, Hes78, dF79f]. Schon [Ros68a]. School [LM61b, Ano92p]. schoolchildren [Mec21a]. Schooling [BF04]. Schools [Cor58]. Schopenhauer [Ein59]. Schou [Kat52b]. Schoultz [Sla88]. Schramm [Fla90]. Schreiber [Rab69b]. Schrödinger [Bur65]. Schubert [Lis58]. Schück [Kat52b]. Schull [Bru57, McC96]. Schultz [Ano86v]. Schumacher [PF75]. Schurr [Lof51]. Schwartz [Ekl80, Sch70b]. Schweitzer [Lib57, Wit95b]. Science [Adv50, Adl57, Adr54, Ahm80, Alf79, Ano45c, Ano45j, Ano46c, Ano47a, Ano47o, Ano47p, Ano50y, Ano50-27, Ano52l, Ano52s, Ano52v, Ano53l, Ano53r, Ano53s, Ano54c, Ano54t, Ano55p, Ano57a, Ano59c, Ano60o, Ano60n, Ano60p, Ano60q, Ano61q, Ano61r, Ano62z, Ano63y, Ano64p, Ano64q, Ano66q, Ano67g, Ano68k, Ano69a, Ano69f, Ano69l, Ano69k, Ano69r, Ano70a, Ano71f, Ano71e, Ano76c, Ano81c, Ano82k, Ano85d, Ano86l, Ano98l, App46, App53, Art68, Ash50, Aug56b, Aug65, Aur71e, Aur71f, Aus60, Bar60, Bar57, Bas68, Bel57, Ben59, Ber53a, Ber53b, Ber49, BE62, Bha71, Bix49, Bla61a, Bla69a, Bla67, Bok49, Bor58b, Bri49, Bro68, Bro82, Bro56c, Bro58, Bro71, Bro05, Bur65, Bur66, Bur57, Cha69b, Che70, Coc52, Com48c, Con46b, Con65]. Science [Coo56, Dad71, Dar51, Dav53, DeW60, Ded65, DG79, Dew52, Dob49, Dob52, Doo59, Dou68b, Dre51, DuB46b, DuB57b, DuB66, DuB69a, Dub55a, Dub56a, Dub56b, Dub50a, Dun48, Dur69, Eas85, Eba61, Ede58, Edm74, Edm75, Eis57, Ell56, Ell71, Eml66, Eps52, Esp72b, Fal68, Fel69b, Fil76, Fin56, FD78, FS84, Fin84, Fin54b, For57, For64, Foz62d, FU63, Fra57, Fre60, Fre68b, Fri56, Fro49, Fuc70, Gal56, Gan52, GA73, Gen84b, Gen86, Gil65, Gil68, Gil72, Glu50, Gol96a, Gol78a, Gol69c, Gom70, Got20, Gra82a, Gra73, Gre67d, Gre07, GS68, Gro52, Gro53c, Gro56a, GC83, Had56, Han69, Has67, Hav04, Haw47, HBM^{+}07, Hei73, Hil52, Hin67, Hoa59, HR51, Hol03, Hor77, Hor66, Hor68, How56, HM72, HUC48, Hux49, Int67, Iff67, Inf62]. Science [Ing56a, Jac74c, Jam58, JOO15, Jan76c, Joh60a, Kap62, Kap94, Kas69, Ker68, Kev07, Kil59, Kil62, Kil65, Kis69, Kis60, Kis72, Kli52, Kom54, Kov68, Kra06, Kui73, Kus68, Lan08, Lar67, Lar63, Leg56, Lei70, Les63, Lew56, Lew61b, Lew67c, Lew70a, Lew71e, Lil49c, Lin69, Liv68, Lon57, Lon73, Lon51, Lon52c, Lop66, Lop68, Mac67, Mac68a, Mac65b, Mac66, Mal50, Man59, Man57, Mar49b, Mar54b, LS59, Ma’52, Mat50, May47a, McD62, McE70, McL66, Mea69, Men68, Men71, Mey49b, Mey56, Mil81, Mor57b, Mor53a, MS77c, Mor79, Mor47a, Mor47b, Mor51, Mor48b, Mot51, Mul52, Mul59, Mul77, Mum54, Nac65, Nef59, Ner69, Neu65, Ney69, Nie65, Nor64, Noy47, Noy57, Old66, Old69, Opp51c, Opp56, Opp59a, Opp59b, Orl84, Orr56b, Pop81, Pop52]. Science [Pre48, Pal02, Pan81, Pan07a, PB68, Par46b, Par47, Pas66, Pas74, Pec70, PD49, Pei53, Pie54, Pie58, Pie66, Pla73, Pol46, Pol49, Pol50a, Pol57b, Pow47, Pri47, dSP65, Rab69a, Rab46f, Rab52f, Rab52i, Rab53c, Rab53g, Rab56g, Rab56k, Rab56l, Rab56j, Rab57h, RS58, Rab59g, Rab59f, Rab59k, Rab61i, Rab62a, Rab62b, Rab64f, Rab65h, Rab65d, Rab66a, Rab68d, Rab69j, Rab70c, Rab72a, RMK65, Rei01, Rei68a, Rei72, Rep89, Ric52, Rid46, Rid47, Rin75, dQR70, Rot72, Rot58b, Rub08, Rus55b, RB65, Rus62b, Stu47, Stu48, Sci49, Sab72, Sac57, Saw79, Sch60b, Sch58a, Sch58b, Sch69c, Cou54, Sea67, Sha93a, She71, She77, She64c, Shi55, Shu72a, Sie52, Sil65, Sim84c, Sim46a, Sim54d, Sin67a, Sin79a, Sin81, Sin56, Sko68]. Science [Smi59b, Smy50a, Spo72, Sta49, Ste69a, Ste46, Ste65a, Ste61, Ste65b, Sto56b, Str71, Str05d, Sut68a, Sut84, Sut16, Swe93, SG69b, SG69c, Tay52, Tel66a, Tel71, TS01, Tel66b, Tél68, Tel73, Tel76, Thi59, Tod64b, Tod78b, Tod80, Tod81, Top63, Tru50, Tsi77d, Ven79, Vis56, Voh71d, Wal69, Wal64, Wat51, Wat63, WW50, Wea57, Bus46, Wei56a, Wei58, Wei66c, Wei69a, Wei59b, Wei65, Whe64, Whi64, Why71, Wie62, Wie64, Wil53c, Wil81b, Wio79a, Woo92, Woo56, Wyl51, Arc54, Zim75, deF51, dF79f, dF79h, dH75, dJ63, dL68, vL63, AB77, Ack89f, Ahe91, AM21, Ano78p, Ano88s, Ano96o, Ano97j, Arn98a, Bla77, Blu88, Bor57b, BHH11, Bri80a, Bru76, Cas92, Cav82, Cha82b, Cha86, Day77m, DR92, Dou15, Eav16, Eav18, Edw12, Ein59, Epp20, EG24]. science [Fel92, FR98, Gar77a, Ham14, Har76b, Hes78, Hsu79, IS15, Key80, Kow77, Kra15, Ksa91, LL90, Lak78, Lan76, Les93b, Lur77, Mac77b, MK79, Mor83a, NR79, Nor88, Par85, Pep15, Pet11, Pol76, Rab85, Ric76, Rog88b, Rot97a, Rot92h, Rot97f, SSK10, Sch20, Sch88b, Sha91c, She97, Shi47d, Shu78, SH89b, Sto19d, Str81, Tho12, Wil87, Zim78, Zim81a, Ano46-53, Ano50-38, Ano50-43, Ano52w, Ano52x, Ano53w, Ano78l, Ano80f, Ano87n, Bar81i, Cra70a, DuB50b, Fer57a, Hei73, Kam74, Kan77a, LS20, Lov66, NK20, Opp20, Pla57b, Rab72g, Cou50, Shi47d, Voh71b, Wie64, dF79f, dF79g, Ano78e, Ano78n, Ano94h, Bal88, Col84b, Cre68, Fay75, Gen84a, Gil85, Gru70b, Kat54a, Mot68, Nor06, Smi75c, Sno68b, Sno69, Spi66a, Tuc72, Wig62a, Ano60f, Ano89w, Bar66a]. Science [Bri80b, Can70, Ded62a, Fel73a, Fer57a, Gol55, Haw56, Haw58a, Her88b, Lan58, Lew62a, M.84, Opp62a, Ost72, Pfe52, Pla57a, Rom90, Rus51a, Sim62, Sim65, Soy89, Tel76, Tod78b, Vuc69, dF80]. Sciences [Ano85a, Ano88e, BS57, Bok48, CLR06, LeR49b, Mul64a, Myr73, Nor64, Opp55, Par47, Que52, Why62, Wig51, Woo61, Ano11e, Gro84b, MH84, Pal84, Ano45m, Ano46-47, Ano72f, Ano86w, Rab61h, Haw58b]. Scienomics [Aur71d]. Scientific [Ada52, Ano45m, Ano46-30, Ano46-108, Ano50u, Ano50z, Ano50-36, Ano50-48, Ano64o, AM74, Apt61, Axe60, Bal82, Boo46, Bou70, CS75, Car49a, Chi69, Com63b, Dub62a, Foz62a, Gab73, Gar66, Ger49, Gla58, Gol96a, Gra83, Gre67a, GLMS60, Gué69, Has60, Hei64, Hum59a, Hyd75, Kat52d, Kat53b, Kat57b, Kol75, Lof60, Lon69b, Low65, Man49b, Mar58d, Mar63, McN63c, Mei51d, Mit70, Mor73, Myr73, Nie55, Nix60b, Nor64, Noy46b, Oro59, Pol50b, Pri53, Rab46d, Rab51d, Rab52h, Rab53d, Rab63g, Rab63h, Rab63f, Rab63i, RC70, Rot62, Sci51, Sci46, Sag73, Sal75b, Saw80, Say55, SD17, Sch84b, Sem53, Shi54a, Shi61, Shi76, Sin65, SD55, Smi74c, Smy51, Sol66, Sta60b, Sul58a, Top62, Upt63, UM67, Wad58, Wal51, Wei52a, Wil59, Wil69c, Zah69, Ano83b, Atk78, BDE^{+}77]. scientific [Bey12a, Cha83a, DM24a, Fie21a, GMV^{+}24, Ham76, Hes85, Kuz88, Lef77a, Odh79, Spr25, Wei76b, WW80, Ano46-76, Lan56b, Pri68, Gen84a, Hyn69, Nad79]. Scientist [All60g, All62b, Ano47d, GS51, Ano54c, Ano54b, Ano72g, Arm53, Ash50, Asi62b, Bat73, Bél48, Ben56, BB63, Bri49, Bro56b, Bro72, Cas77f, Cor53, DS49, Day75o, Dub56b, Ein52a, Fel70a, Fel92, Foz62g, Fra47, Gol75, Gro50, Had56, Hay64, Hei70, Hig66, Hoc57, Hod59, KK72, Kra46, Kus68, Lew71b, Lof60, Mal65, Mar49a, Mar48a, Mas48b, Mic51, Mor57a, Nad72, Nis47, Ogb48, Opp51c, Pas55a, Pei50, Pei69, RE68, Rub08, Rus60, San66b, Sch00e, Sha63, She64a, Shi57, Sie72, Sim46b, Sim60b, Smi75b, Spi49, Swa49, Ure46a, Ure48c, Wie48, Zim48, vL48b, Ano90-61, Ano10c, Ano15b, Bes85, BFG^{+}85, Dro22i, Ein79b, GB88, Gor76, Gri83b, Hoa77, Hor77, Lop76, Mar79c, Med81, Smi65, Smi70b, Sto17b, The90, Wie57, vH87, Ano47h, Ano55m, Ano55v]. Scientist [Ano56k, Ano78o, Mei73, Ack88a, Dow79a, Gor71, Har56, Key81, Moo94a, Sch54a, dF79h, Ano54k, Rab48c]. Scientists [Ack84, ACU^{+}54, AAF^{+}52, ADH^{+}73, Ano45b, Ano45n, Ano46-27, Ano46-49, Ano46-65, Ano46-89, Ano46-88, Ano48f, Ano48v, Ano49l, Ano49-33, Ano49-35, Ano51-29, Ano53-40, Ano53-39, Ano53-46, Ano54-28, Ano54-27, Ano57k, Ano58l, Ano58k, Ano62-27, Ano65g, Ano66j, Ano66r, Ano67j, Ano67o, Ano68h, Ano68l, Ano69a, Ano75g, Ano78i, Ano14a, Aud65, Bak75, Bar57, Bea47b, BAB^{+}46, BS47, BAA^{+}75, BWH80, Bri48a, BEI^{+}55, BCH^{+}48, BFM^{+}49, Bur50, Bur74, Bur62b, Cah76, Cas77d, Cha78a, Cha70a, Cha69a, Coa88, Com57c, CCMU46, CFU47, CD48, Con48, Con52, Con58, Cou71, Cra68, Dam61, DP53, Dan54b, Day75c, Der88, Dow83d, Dub59, Eme47, Ele57, Eid66, Ein48c, EBU^{+}52, Eis80, Etz64, Fed68a, Fed68b, Fin57c, Fri56, Fri70a, Gar83a, Gla62c, Gol93b, Gol95b, Gol82a, Gol78a, Gou47a, Gra86]. Scientists [GW65, GR63, Gru70a, Hal56, Hig50, Hil46, Hin67, Hir55, Hol93, JCKG^{+}50, Kyo85, Kas65, Kat57b, Ken71, Kor92, Kra15, Kuz04, Lan90n, Lap59b, LT70, Ley51, Lof47, Lon67, Lon58, McC65, Mei48, Men69, Mia15a, Mil47, Mit62, Mor66a, Mor66b, NB78, OW65, Pas51, POF49, Pau57, PR62, Pau67, Pip55, Pow59, PvH72, Rab46e, Rab47f, Rab49a, Rab51e, Rab52i, Rab53h, Rab56e, Rab56g, Rab57h, Rab59j, Rab65c, Rab69i, Rab69k, Rab71d, Rid51, Rin75, Rit48, Rus56a, Sch63b, Sch69d, Sch54d, Sea56, Shi54d, Sim47, Smi82b, Spo62, Sta48, Ste62, Ste75a, Sti62, Sut70, Swa66, Swe87b, STS^{+}49, Tel46a, Tel47a, Tel63, Til60, Tol58, Toy84b, Tru48b, Tsi80, Ure48b, VFIS48, Wea79, Wei54b, Wie47a, Wig62b, YTS61, vL48c, Ack89g, AO98, Ano85y]. scientists [Bar81f, Bar81i, Boy85, Dw77, Dis79, Dje76, Dow92, ET20, Fre77b, Fra76a, Gab86, Gol86b, Har10, HM15, HRS81, Hug87, Hun85b, KW15, Kap92, Lem81, Lop78, Mal85, Man15, Pug76, Pug78, Paj52, Pru85, Sag93, Sch15b, Sch75e, Shi47d, Spe95b, Tob12, Vil75, Wil80d, Wil85b, YG77a, AM86a, And10, Ano45g, Ano45h, Ano45m, Ano46e, Ano46-33, Ano46-35, Ano46-42, Ano46-83, Ano46-89, Ano47k, Ano48-35, Ano49k, Ano51i, Ano51y, GS51, Ano52y, Ano68l, Ano69d, Ano69k, Gro50, GR63, Hig48b, Hil51, Kil57, Lev51a, Nel75, Pla59a, Qua80, RG49, dF79f, dF79e, DeW81, dF79e, Ano49j, Ano50g, Ano51d, Ano52e, Ano53j, Ano54f, Ano55d, Ano56e, Ano57e, Ano58b, Ano59b, Ano60g, Ano61e, Ano62m, Ano64h, Ano65d, Ano66a, Ano67c, Ano68b, Ano69b, Ano71a].

Scientists [Ano72d, Ano73a, Ano74a, Ano75b, Ano76a, Ano77k, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano78h, Ano78a, Ano78b, Ano79b, Ano80e, Ano81d, Ano82b, Ano83e, Ano84c, Ano85e, Ano86m, Ano87t, Ano88f, Ano89h, Ano90q, Ano91t, Ano92k, Ano93o, Ano94i, Ano95h, Ano96h, Ano97h, Ano98i, Day74a, Gou78b, Hur78a, KW15, Ken62, Mec20e, Mes64b, RG49, Sch69d]. Score [CMMF62]. scoreboard [Jac77]. Scorpion [And71]. Scotia [Can79]. Scott [Boy05, Cra83b, Gol92a, Kin86, LP92, Mos92, Pon67a]. Scoville [Ano85v, Wal81a]. scram [Hor88a, The88]. scrambles [Ano91-29]. Scrambling [RM16, Ano91-29]. Scrapping [Joh94e]. scraps [Woo96a]. Scrapyard [Gre73b]. Screen [Sul53]. Scribner [Ano86k]. script [Man80]. Scud [Bul02e]. Scurry [McE52]. SDI [Bre87, Ano86d, Ano88g, Ark84d, Cha88b, Cla91, Coh86, Dar86, Dre87, Fet87, Gar87, Gar86, Gol86b, Gur88b, Isa89c, Isa91m, Kre88e, Mar85b, Mar87a, Par88a, Ros88, Sam86, Sap88, SLM87, Ste87, Van86, Wig88, Woo92]. Sea [All62a, Ano87i, Ano87-39, Ano90p, Ark87a, Ark87b, Bar87, Bre74, BKSG60, BR15, Fie87, Han87a, Hol16, Jac15, Jon65, Lew64c, Mcc87, NA89l, NA91k, Opp03b, Pon65, Pon67b, Pos87, Ros75, Sca11, Sti18, Van20, Wal61, Zen76, Ano88p, Ano89u, Ark83b, Ark85g, AH89b, Che97, Dro24c, Hud77, McL18, Mil79, MHKL18, NA87k, NA88j, NA90i, NK16, Nuc18, Tre18, TM98, Ark88b, Coc89, Law11, LL16, Pat75, Rev83, Swi83, You78a, Jas81]. Sea-Based [Ros75]. sea-launched [Ano88p]. sea-level [Dro24c, McL18, Tre18]. Seabed [Alb73, McK74]. Seaborg [Lon82, Lew72c, Moo99e]. Seabrook [Jas81, Swe78a, Abb91]. Seal [Myr75]. Sealed [DM05, Pol62b]. Sean [Kri85]. Search [Ano48n, Ano81f, Bak75, CLR74, Chi57, Gen84c, Gus85, Hor71, Joh99a, Kov68, Moo95b, O’C87, Rea68, Som84c, Abb82, Ano91-43, Bar83a, Ber77b, Bro86d, Jon84, Lan16, iM16, Mor87a, She16, Tay21, WW86, Pla59a, Ano85d]. Searched [Mos67]. Searching [Ano93s, Fly02d, Rot96k]. seas [Kop15, Nas18]. Season [Alv01a, Bou97, Coh92, Cre90]. Seasons [Ame68, Kin78b]. Seat [Lan08, Dal92b]. Seats [Isa02d]. Seattle [Gal81a]. Seco [Knu89]. Second [Ano47s, Ano15f, Han75, Hen71a, Kap91a, Ken79, Lap59e, Leg56, Lew73b, Mil52, OJCP^{+}47, Rab66e, SNB^{+}54, Wri63a, Bar80k, Day75a, Don89, Eps81, Gar83a, Hib94e, Lie46, NA88o, Pug80, Ray90, Adi69, Ano46-90, Ano58g, Coc59, Dow81g]. Second-generation [Ano15f]. second-strike [Ray90]. seconds [Lev96]. Secrecy [Adl85, Aft00, Ano46-44, Ano49-36, Ano51w, Ano51z, Ark85i, Ber55, Cha82b, Che82, Coc81b, DM81b, Fel76g, Fin55, Gla75a, Gui65, Haf05, Hal85, LB53, Lap56c, Les81, Lil53, Low76, Mar85a, McD08, McE88, Mos61, Nat85a, Ret99a, Rid46, Shi56b, Wal52, Ano85o, Ano86-32, Car83a, Cha86, Kal82d, Kal89, Lan83b, Lan92b, Man84c, Zac76, vL92, dKL83]. Secret [Ano66f, Ano93-35, Ano95i, Chi59b, Dow86, Gol88a, Gre82b, Han92b, Har88c, Hig48a, HM90a, HLL03, Hug56, Kla88, Kre89c, Lev84b, Mar59a, Moo95d, Mos90, NA87j, Nor04, Roc76, Sel59, Shi54c, Swe93, Ber87c, Day79, DeW79, Eav18, Fel79i, FS95, Fre91d, Ham98, Hes97, Isa91l, Kno90, NCA88b, NA89i, Pow93, Gre82b, Kor92, Tay98]. Secretary [Ali57, Bro87b, McN67b, Mec19c, Ano46-91, Ano50-29, Ano58i, DP12, Dro16a, Dro23a, For49, FHN^{+}46, FHN^{+}63, McN67a, Per20]. Secretary-General [DP12]. Secrets [Adl85, Alb00, Ano49-27, Ano50s, Ano03g, Pai82h, Rab47h, Rot95g, Rot98a, Rot99b, Sto99, WH00, AK98a, Ano87-35, Ano91-37, Ano93u, Ano94l, Ark89a, Bar81g, But97b, Che52, Dan90c, Gea89, Han89, Han95c, HM97, Joh97a, Klo86, NA93o, NA93s, Rot92b, Sag93, War86]. section [Ano83n, Ano45c, Ano45j, Gol21]. Sections [Ano49d]. sector [CR17, Sam22]. Sectors [HAIT80]. Secular [Kee07, Sta49]. secure [Ada82]. Securing [MD24, She56, Ano87l]. Security [Adv60b, Ano48a, Ano49b, Ano49-37, Ano53y, Ano54b, Ano55a, Ano55i, Ano55j, Ano55v, Ano55w, Ano74i, Ano82i, Ano88e, Ano89d, Ano93m, Aro55, Aug48c, Aus47, Bec55, Ber46, BFW47, Bon53b, Bro55c, Car89, Cav47, Cha83a, Chi57, Com48c, Con55, Cus49, Div79, Dow85, Dre82b, DuB55b, Ein50, Eis57, FT76, Fel76g, Fel77o, Gla56a, Gor51b, Gou55, GM54, Gre55b, Gre58, Gre64b, Gre05, Gro56b, Gué66, Her83a, Het55, Hig46, Hil60, Hoo81, Ing56a, Isa00c, Isa01c, Jac69, Jac80b, Kal82d, Khr05, Kib12, Lal64c, Led66b, Leg57, Liv56, Loe02, McN64, Mor55, Mot48, Nay12, NA01, Pug71, Pub55, Pal55, PBG55, Pit64, Pri07, Rab54f, Rab54i, Rep86, RD76, Rot02e, Rub08, Sal64b, Saw54b, Sch08a, Sch05a, Sch04d, Los54, Sha06, Sha98c, She83b, Shi55]. Security [Sie04a, Sig85b, Sim54d, Sim54e, Sla88, Soh64, Som84c, Squ12, Ste07b, Ste08, Swa49, Szi54, Tsi76, Uhl98b, Wes00b, Win71, Zac50, Ack88b, Alb83, Ano76b, Ano85s, Ano86-27, Ano86x, Ark83a, BOW10, BBL13, BB79, Bur88, Bye91, BB22, Cha82c, Cha85a, CG12, Coh92, Deu91, Dum76, Eps82b, Eva54, Evr78, Fai81, Fel81g, Fel82i, FC88, Gle91, Gle94, Hal85, Her79, Isa93e, Joh91, KK12, Kok88a, Kos10, Kri14, Lan90l, Lei83b, LC78, LL16, Lin79, Mak90, McC85, Mec16d, Mec22a, Moo95n, Mye87b, NF13, Nat88, O’B79, Ode88, Pug76, Pug77, Pug86, Pai82h, Pla84, Por19, Sch21, Sco08b, Sha87d, Squ13, Sta81, Ste89a, Ste85c, SG69a, SG71, Sto24, Sub77, The89f, Toy84a, Toy84b, Tsi87, Tur13, Väy87, VU22, Wei78c, Wei78d, Whi79, Woo12, Yaz88, Yor75b]. security [Yos80, Ano47f, Ano47j, Ano47s, Ano55m, Ano56k, Ano83b, Dav48, Mye87a, Sei60, Eis84, Ano89g, Bou87, Gar72, Gol95d, Hun87, Lap61a, Moo97b, Nel84, Phe56, Sar78, Swe89a, Tir88, Ano89w, Don86a]. Seduced [Gui07]. See [Inf65, Ura01, Ano98p, Fel78d, NA94e, Rot92p]. Seed [Smi97]. Seeing [Ano12b, Isa90p, Kim06, Bra13a, RK93]. Seek [Pom65, AH91e, Isa92h, Kes89]. Seekers [Str07d]. Seeking [Ano91-39, Ano11e, Ege16, Ike16, LM71, Ric00a, Sid16, Sto79]. Seeks [Gol77, Smi75a]. Seen [Hel70, Wei46, Dor85b, Rot92h]. sees [Ark95d]. Segal [Sim85, Ive16]. Segrè [Wil71, Kat54c]. Sein [For80]. seine [For80]. Seismic [Lei83a, Wil83]. Seismographs [vMC98]. Seismology [Lat73]. seize [Hib95e]. seizes [Isa91d]. Seizing [Ber93b]. Selected [AM51, Ano50-44, Ano62p, For80, Sch49, Ano87-31]. Selecting [Lwo60].

Selection [Dud77a, Ano69d, Ano97j, BM22]. selections [AM86a, Ano47k]. Selective [Ahm03b, ACC^{+}67, Ano96m, Sco94b, Ano50-48, Cro68, Pub68]. Self [Dw77, Kli81, Ric75b, RT12, Rod69, Wei61, Wel60, Aye85, Cle89, Dro22f, Eng78, Fre76a, Hoa77, KDFS81, MvE13, Nor04, Par85, Ano78l]. Self-assured [RT12]. self-censorship [Par85]. self-contained [Eng78]. self-destruct [Cle89]. Self-Destruction [Wel60, Hoa77]. Self-Destructive [Ric75b]. Self-Enforcing [Wei61]. self-fulfilling [Fre76a]. self-governance [MvE13]. Self-help [Dw77]. self-published [Nor04]. Self-Reliance [Kli81, Ano78l]. Self-Restraint [Rod69]. self-sufficiency [Aye85, KDFS81]. Selig [Bro47c]. Selin [Car94b]. Sell [AM74, Cah75, CLD88, Far08, Ano92r, RM16, Sez16]. sell-off [Sez16]. Sellafield [Opp03b]. Sellars [Bul08n]. seller [Bar81c]. Selling [Ric74, Wes76, Ano94o, Isa90h, Isa90m, Nor88]. Semantics [Fel61, Ram97]. Semi [Ano48-28, Ano48-29, Ano98o]. Semi-Annual [Ano48-28, Ano48-29]. Semi-super [Ano98o]. Semiannual [Ano49d, Ano49c, Bos51, LeR49b, Tur50, Tur51, Tur52, Tur53]. semiconductors [Cha89c]. Seminar [Ano96v, Ano96w, AC65]. Semipalatinsk [War95]. Semler [Ano87s]. Senate [dF79h, Ano45e, Ano45l, Ano45o, Ano46-31, Ano46-32, Ano46-55, Ano46-61, Ano46-82, Ano46-92, Ano47c, Ano47t, Ano47u, Ano47v, Ano86v, Hum59b, Isa87b, Isa87e, Isa87d, Isa88c, Isa90q, Isa92f, Isa92h, Isa92n, Isa95c, Isa95d, Isa96c, Isa97e, Isa98a, Isa98d, Isa98f, Isa02b, Leb79, Men91c, Nel09, NK86b, Sch79c, Smi98]. Senator [Gol67a, Rod70, Sch71a, Ano46-93, Ano47x, Ano47w, Ano48n, Ano48-37, Ano50-45, Cla66, Con48, Fox52, Ken71, Opp49a, Pai85b, RDS^{+}52, Rab52g, Sim59f, Van52]. Senators [Ano51z]. Send [Han92a, Ano93z]. Senghaas [RK87a]. Senior [All14b]. Sense [Ano95p, Ano95v, Aro55, Gro78a, Haw58b, Isa90t, Isa02c, Kat54a, Mul77, AGB22, Mat83b, Rob91a, Rot92b, Rot98d, ZGD83, BWH80]. Senseless [PY98]. Sensibar [Ano88j]. Sensibility [Ker76]. Senso [Ano91-47]. Sentenced [Gla75a, Ano78q, Dow80i]. Sentinel [Sel69b]. Separate [Ste75b]. Separation [BF05b, Day75n, Fel66c, Gar77c, GC77, BDF^{+}92, Cas77e, PB84]. separatist [Coh92]. September [Ano58l, Ano58k, Ano63o, Ano64q, Ano68l, Ano84d, Haf54, Int67, Rab67h, Ano92-27, Cam53, McN67a, NA91a, NA91c, Nor94a, NA95b, NA96b, Rot01a, Woh78]. sepulcher [Gol11]. sequel [Hec19]. sequences [SSK10]. Sequencing [Lin08]. Serb [Kri95]. Serengeti [Man15]. Sergei [Bou98b, Leb91b, Kap90a]. Sergio [Rez07b, Bul09b]. Series [Ano53e, Ash81, Dow82h, Han98a, KA90, Rot94j]. Serious [Sch06b, BKV13, Lev84a, Sut91]. Sermons [Bri86]. Servan [Rab69b]. Servan-Schreiber [Rab69b]. Servant [Myr69, Sta58]. Servants [dF79a]. Serve [Loo50, Jen74]. Served [Fly00e]. Service [ACC^{+}67, Ano50-48, Rab61i, Sci55a, Sap68, Wat63, dH75, Brz89a, Cou77, Dan54a, Ste84, Tsi77a, Bri50, Cro68, Gar68, Pub68, Dam61]. Services [Sal53, Ded78, Wie47b]. Session [Ano78m, Hud78, Noy47, Sha78a, Sha78b, Ano80p, Her88a, Mat82b, PC92a, Fel78h, Wei78a]. Set [Ano96l, Day74e, Squ12, Ack88d, Ano77l, Car03, Dav55, Fig23, NA90l, RL78a].

Seth [Gov93]. sets [Cav13]. Setting [Fel80h, Lor01c, Sch67a]. Settle [Man64, RK89b]. settled [Loc91b]. Settlement [Chu50, LC78]. settlements [Mei77a]. Settles [FvH99]. Seven [Ano69s, Ano02j, Gai69, Hur88, McC88b, Mer47, Dow82i, Tsi82d, Yor82]. Seventeen [Jac64c]. Seventh [Tur50, Wag61, Lee24, Rap11]. Seventies [Sin70]. Several [Wil71, Kis76a]. severe [Cro17d]. Sewage [Jac74b, Mar71b]. Sewerless [Lei75]. Seweryn [Fis83]. Sex [Ano97j, SHT^{+}87]. Sexist [Hob78]. Sexless [Hec78]. Seymour [Cav58, Cro65b, Sid69, War92a]. Seyom [Bal79c]. Seyon [Mes84b]. Seyyed [Ano15g]. Shabby [Gil65, Sim65]. Shackling [Moo96j]. Shades [Sch90, Arn92b]. Shadow [Opp68, Say84, Thi55, Uhl95b, Mar93a, Mia15a]. Shadows [Ano81k, Ric05, Tur17, Veg99, LS94, Low93, Dow81b, Tum87]. Shah [Bre93]. Shaheen [BD79]. Shake [Mil62]. shaken [Isa98f]. shakes [Lar89]. shakeup [BMPd89]. shale [Lev12]. Shall [Ano53-48, Lev47, Lev48, Szi49a, Szi49b, Szi63a, Wei53b, Wri77]. Shallow [Dav76a, Dav77b]. shalt [Eps90a]. Shame [Edm77, Ano92y, Har22]. shaming [Dro22f]. Shanor [Klo85]. Shanto [Oze92]. Shape [Deu83b, Pal64a, Fer88]. Shapes [Pow90a]. shaping [YG77a]. Shapiro [Met83, Gar83a]. Shaplen [DeB81a]. Shapley [Ano86l, Mea93, Dra69]. Shapo [Tho80g]. Share [AT92, Jac74b, JPR90]. Shares [Rab48h]. Sharing [But97b, LB53, Mor83a, KKJK23b]. Sharma [dF79g]. Sharp [Ano78m, Sta85, Tay87]. Sharpen [Gom00]. Shattered [Fel77o]. Shatterer [Ber75b]. Shattering [Swe86]. Shaw [M.85]. Shedding [Rot98h]. Sheen [Ano46-107]. Sheep [Ano96m, MD75, Ano95i, Dri93]. sheer [Rob91a]. Sheets [Mue77]. Sheila [Ahe91]. Shelanski [Sti83]. Sheldon [Fre95, Ing70, Lon78]. Shelf [Pri07]. shell [Pai85b]. shells [Die91]. Shelter [Ano62-28, Ano03h, Dee57, Fed62, Hus53, Rot89a, Ste57b, Ano56c]. Shelters [AS62, Aue06a, Cha63a, Dys62, Dys63c, Fel62c, Foz61e, Ing62, Lap58, Mor62a, Sin61]. Shenfield [Par88b]. Shep [Dow81g]. Shepherd [Sto19d]. Sheriff [Sut04]. Sherr [Lep89]. Sherry [Jer87, Uhl95b]. Sherwin [Cah76]. Shevardnadze [Sho92b]. Shield [CGO05, Gor59, Gre71, Hir88, Dro19, WB87]. Shields [Has57, Wal81a]. shift [Bro16, Fox88, Nag81, Sig87a, Smi21, Dow81g]. Shifting [Isa91n, Ric09, Van90b]. shifts [AH92l]. Shigeru [McK74, Sak69]. Shils [Ano95q, Rei74, Pee78]. shine [Ano95n]. Shines [Lew64b]. Ship [Ano49-41, Ano55g, NA90c]. shipments [BP90, Ial18]. ships [Ano92-29, NA87b]. shirt [Ano93u]. Shlomo [Ram88b]. Shock [Lew76, Moo02d, Dic01, Dic03, Dic11, Fel77a, Han92c, OW90a, Sha90, Ste77c, Ste60c]. shoe [Alb98a]. shoo [Isa91p]. shoo-ins [Isa91p]. Shoolery [Pal02]. shoot [Rot93u, Swe87b]. Shootin [Ric00b]. Shooting [O’N89b]. Shop [Mit70, AH92f]. shop-till-you-drop [AH92f]. Shopping [ZK94, AH91k, NA88o]. Shore [Woo00f]. Shoreham [Ano72g, Fes73, Sai72, Sch72a, Wil92d]. Shores [Sch68f, Rob90b]. Shoring [War89]. Short [Isa90r, NA88o, Wou81, Ark87f, Fie24, LGW92, Mag88b, McA84, Ram13b, Sim11, Woo98d, Kus83]. Short-range [NA88o]. Shortage [Ber77a, Fis76, Kat57b, Smi74b, TN90, Kat57b]. shortages [Ben76a, Ben76b]. shortchanging [DeG84]. Shorter [Dro20b]. Shot [Jac72a, Tsi75a, Sch76a]. Shots [Jac74b, RD77, Alb98c, Isa91i]. Should [And60, Cle48c, DRM^{+}09, Eme61, Her84, Hum60, Ler53, Lon75, McM49b, Mei48, Mia16, Mor47b, Nel09, RL89, Rho07, Ros68b, Sim84e, Smi54c, Ure48b, Ure50, vW57, Ano15c, AB18, Bel20b, CW23, Dav55, Dis79, Dor90, FTvHW76a, Fie21b, Gol21, Kas21, Kor19, Mec19h, Mec20h, Mec22d, Mei98, Pat90b, Rev21, Rho15, RL14, RE19, Rot77, Sch15b, Sch21, Sol90, Squ21b, SD10b, TS22, Tsi88, Wu23, YSW21, You92, Zik21, vV23]. Shovels [Dia83]. Show [Ell76, Wei52a, Hil16, Hil20, Isa90y, Plo89, Van86, Kur79, Sha94]. Showdown [AH04b, Gun04, Isa86e, Mag88b, Yat69]. shows [Kun23]. shrinking [Squ17a]. Shroyer [Tay98]. Shubik [Kar80]. Shuffles [Cla74]. Shulman [Gov93]. Shultz [Ano13a]. ‘Shuman [Pai81a]. Shunji [Ano90p]. shuns [Flo89a]. Shush [Ano93-27]. Shut [Bro93, Bre13, Ial18]. shutdown [RH12, Sic96b, Wil92b]. Shute [Doh58]. shuts [WM78]. Shuttle [O’L73]. si [Rab61d, Hug95]. Siberia [She71]. Siblings [Hla65]. sic [All59l]. sich [For80]. Sicherman [Ste76a]. Sickly [Kap94]. Sickness [AME50, Ano55u, Bid51, Cle07, Lis46a]. Side [Ant70, For80, Lep89, RKR55, RK55, Ano95k, Han92b, Mil95, OGFM18, Soa96]. sidebar [Par88a, Cas77e, Cas77f, FTvHW76b, Fin77b, Kan77b, Lef77b, Pri75c, Sch75d]. Sidedness [BPR52]. Sidelights [Smi55]. Sides [Gil83]. sideshow [Ade91]. Sidney [Day77b, Dow80f, Dug91, Fre84a, Van87]. Siegbahn [Kat52b]. Siege [Gus99, Jan76c, McC93]. Siegfried [Ano13c, Mec20g]. Siemens [Hib95d]. Sierra [Jag98]. Sieve [Rub74]. sifts [Wur91b]. Sigal [Van88]. sight [Moo95r, Sie07a]. sighted [Fel77b]. Signal [Kap74]. Signals [Gai69, Haf67, Ric00a, Sch65a, Vil75, Wri65, Fie91, Kre86b]. Signed [Spi59, Ano90-55]. significance [BC16]. Significant [Pre60, Ris75, Hol78, Ano98p]. Signs [Ano49e, Sig87a, Sut96b]. Silard [Low93]. Silber [Dow80a]. silence [Day97, Fas78, Isa89k, Kau82, Röl80, TM97]. silenced [Mos79]. Silent [AH92d, Fly03, Gra85, HB10, RK89a, Sid69, Cas76, Fre93b, Ehr79]. Silicon [Dro24d, Kam14]. Silk [NA91m]. Sills [Sti83]. silos [RC87]. Silver [Ano75h, Ano91-37, Ark01c, Sch76b, NA92e]. Sim [Ano97i]. Sime [Gol96b]. Simon [Ano85d, EE85, Rab56h, Sei81, Str67, Dal82]. Simoni [Lyt79]. simple [Sim11]. Simpson [Ano93-29, Dro21b, Moo00g]. Sims [Bra89]. simulations [Ano89o]. Sin [dF79f, Nel92c, Röl80, Rot93s]. Sinai [Day75e]. Since [Ano63m, Bou74, Lal64h, Opp62a, Sca91, Smi72a, Tow66, Uhl95b, Ano89c, Ano07g, PN09, Sch98, Str62, Toy85, Wri63a]. Singapore [Bha71]. Singer [Dow80j, Ras63]. Singing [Ano97j]. Single [Smi74d, Rot92h]. Sink [Car77, Moo97b, Isa85f]. Sinking [Ano92-29]. sinks [Ano91-33]. Sino [Hsi65, Mac65a]. SIOP [Ark83e, Ark00f, Lor01a, Moo97b, NA87i]. SIOP-6 [Ark83e]. SIPRI [Ano86l, G.84b]. Sir [Ber66a, Coh69, Fer57a, Lak66, Lan98, Ler80a, Lit54, Nic54a, Tuc72, Wig62a, Iri56, Pri55, Rab56h, WW62]. siren [ER90]. SIS [Pau88]. Sister [Blo05]. Sisyphus [Gan84, Rab51f]. Sit [Hun03]. sitcom [Boy89a]. Site [Lay07, Ano95n, BM22, CSF95, Gas93a, GK87, Moo97c, NA90k, Spa96, Ano03i, Gea89, Gra93, Tay91]. sites [Ano79l]. Siting [Kea60, Nel74b]. Sitting [Rot95h]. Situation [Bet75a, Fed72, KASN52, Osb49a, Ure46a, Wei52b, Wil67c, Mec22a, Wyl56]. Situations [Wyl56, GT14]. Sivachev [Ekl80]. Sivard [Ano90p, Goo86, Rep83]. Six [Ack88h, Ano49x, Bet69b, Bet77b, Dys69d, Ful69, NA89f, Rat69, Los54, Sok16, Ste69f, Yor69, Ark96e, DH90, McC86b, Rea80]. six-hundred [Ark96e]. Six-inch [NA89f]. Six-Point [Los54]. Six-Power [Ano49x]. Sixteenth [Rab67h]. Sixth [LeR49b, Ano61i, Ney73, Hyd75]. Sixties [Pol61, You68a]. Sixtieth [Opp79a]. Size [McM49b, Pot97]. Sizewell [Pat84]. Skeleton [Aue06c]. Skeletons [Har72b]. Skeptic [Kra89, Kre90]. skeptical [Hui87, Mei20, Sak88]. skepticism [Wea11]. Sketch [Gro56a]. ski [NA93f]. Skies [Mar02b, Spe89a, ZF03, Ack89f, Ano93s, Ano94m, Flo90, Lin88, NA89s, Shr23, Str88, Smi92c, Dia84]. skills [Bun89, Smi83]. skirmish [Bet98]. Skirting [Ano87-38]. Skittish [Gol94]. Sky [Arp61, Bam00, Ben07, O’L73, Sie06e, ZvH01, Ano98l, Ark84d, Ark94e, Bal87, Gor96d, Isa89i, Rot96j, Lyn78]. sky-is-still-falling [Ark94e]. Skylab [Gol74]. Slack [Wri99a]. slam [Dan91]. slant [DH86]. Slash [Jon90a]. Slave [Ano74b, Jan73]. Slaves [Bal84b, Ezr72, Mac84a]. SLCMs [LP91, Mag88b]. Sleep [Thi55, Han93]. sleeps [Alb94e]. Sleepwalkers [Sto59a]. Sleepwalking [Rau91]. Sleight [Ark84f]. Slessor [Nic54a]. Sleuthing [GM93]. Sleuths [Rab53h]. Slick [Coh87, Sto59b]. Slide [Sch67b]. slip [Isa90w, Sei81]. Slipcovering [Gas93a]. Slippery [Shi54b, Sin05a, Mag88b]. Slipping [Fel76i, Gal90a, Gal90b]. slogan [KE17]. Slope [Shi54b, Mag88b]. Sloss [Kin86]. Slotin [Ano46-70]. Slow [Abb04b, Alb93a, Ano91-34, Pit52a, Pit52b, Rab46c, Sch04d, Fel76d, NA88k, NA89a, Roy76, vHJ97b, Fie21a, Pra03b]. Slow-Walked [Pra03b]. slowdowns [Loc91a]. slowed [Moo25]. Slowing [Ben02, KN14d, Sco77]. Slowly [Ing77, WC17]. slows [Mil12a]. Sly [Mei90]. Small [Bél48, Ber72a, CJLT24, Dav07, Ela93, Fin57c, Gab01, Gou61, Kih47f, Kos04a, LH91, Moo99h, Not15, RR07, Sas06, Sch07d, Ven80, Wad74, Ano90-31, Ark84c, Arn98a, Fei90, MR21, Tha63, Usm79, Zam95a, Zha12, Har79b]. smaller [Mar84]. Smart [Ahm04, Ano91-40, Ark00g, Bel93, Cha89a, Dro22f, Fie18]. Smarter [KK98b, WP19, Gei17, Kre95]. Smidovich [WS21]. Smil [Mye93b]. Smile [Bri05]. Smiling [Ano96m, Ara94b]. Smith [Ano78n, Hig66, Kre90, Shu78, Ano91s, Bro54, Els90, Kre83a, Lap69, Ray81, Rez07a]. Smithsonian [Ano98p, Gol95c]. Smoke [Bla84, Weg96b, Sul53]. smokes [Rot93k]. smokescreen [SP82, ZM96]. smooth [Law90]. Smuggler [Hib06]. smugglers [Spe86]. Smuggling [Lee97, Hib95e, Hib95f, Spe87b, Weg96b]. Smyth [Kra78]. Smythe [Fal61b, Fal61b]. Snafu [Gar91]. Snapshots [MK03]. snare [Hol84b]. Snatching [Isa91o, Nel92b]. sneakers [NA89c]. sneaks [Ark89d]. Sneha [Dro23d]. Sniffing [Mag88b]. snooping [Ano90-45]. snoops [Ano96j]. snoozes [Isa97d]. Snow [Pla59a, Dro20b, Dro24g, CMMF62, Fal61a, Lan55, Mar72a, Sim63, Pow89b, The89e, Smi54b]. Snubbed [Kap94]. snubs [Isa95c]. Snyder [Wol63]. So-Called [Szi60a, KGS25]. Soal [Ras55]. soap [Ano92t, Mar89c, All59l]. Sobell [Nat67]. Sochi [Rab70c]. Social [Ano69f, Ano78p, Ano90p, Aug56b, Bel57, Bla69a, Boy85, Bri48a, Cog82, Com57c, Dev69, DuB69a, Dub56b, Eri95, Fla65, Foz62g, Fra47, Goo86, Gor71, Gra73, Gro56a, Had56, Han66, Hor66, Hor68, Hut53a, Kap66, Kat53a, Lea70, Mac61, MH84, Mat50, Maz69, Mei69, Men71, Mor51, Mum54, Myr73, Ogb48, Olu66, Ost72, Par47, Pla57b, Pla61, Pla64b, Pol67, Pow82, Rab56g, Rab75, Rep85a, Rub73, Rus60, Sch68c, SK69, Sno69, Spi72a, Ste61, UM67, Wal69, Wei66b, Wil71, Zag92, dF79h, Ano93y, Ano11e, Dro22f, Kal77a, Kas12, Mei77b, Orl84, PO16, Rog88b, SH89b]. Socialism [Bir76, Klo85]. Socialist [Kap91a, NM49, Van90b]. Sociality [Saw79]. Socially [Sie72]. societal [Ger13, Ibr13, Zer13]. Societies [Ano45m, Ano69l, Nad72, MS77b, Rab73c, Spi72a]. Society [Ano48e, Ano71e, Ano76c, Ano87n, Aro64, Bar79b, Bas68, Bat69, Bet75b, Bri58, Bro75, Cha70a, Dar51, Edm75, Has60, Hoa59, Kel79, Lei75, Lev51a, Mas48b, Mot51, Myr66, Myr74, Nor64, Ost72, Pol50b, Sch68f, Shu72b, Ske90, Smy50a, Szi61a, Top63, Tri71, Ure48c, Wie48, vL48b, Bro79d, Bro79e, Dis79, Kra15, Sto19d, SCHD20, Wie85, Alb54, Ano45m, Ano46-27, Ano46-88, Ano65g, Ble54, Hai54, JC54, Kat54b, Mer54a, Rab54c, Rum54, Shi47d, Tod55, Wei69a, Zim75, Ehr83, Sco71, Vuc69, PF75]. Sociobiology [Guh65, Sco66]. Sociology [Ano78n, Ano78l, Pit91, Smi75c]. Sociotechnical [Tra24]. sodium [PR14]. Sofia [Pav98b]. Soft [Aft94b, Buk93, Dow83c, Low76, Mac86b, MA17, Myr69, Sch68d, Way65, dV78, Cav92, Fel77e, Rot94h, Nef78]. soft-headed [Fel77e]. Soft-kill [Aft94b]. softens [ZK94]. softly [Gey94]. software [Mye86, Coh86]. Sohlman [Kat52b]. Sohn [Sha59]. Soil [Gya90, Kel64, Sto19a, BMPH18]. Soil-Use [Kel64]. Sol [Axt73]. Sola [Ano78p]. Solar [Ano78d, Bro76b, Cap77, Cra76, FK71, GB73, Gru78, Hou76, Hug95, JS79a, Kam14, Kel79, Kli81, Mak76, Maz81, MM71, Mer72, Par73, Pit67, Rom77, Ste73a, Sto72, Tay81, Tel51a, Zen76, vHW75, Ano11h, Ant76, DJ13, GJS80, Hir76, Shu76, Wol76, Woo12, vHW77, Pat81a]. Soldier [Har07]. Soldiers [Bur74, Hue77, MM01, All14b, NA89c, Wes97]. solitary [Rot98g]. Solla [Opp62a]. Solly [Her90a, Kil78, Lak66, Tuc72]. solo [Ark98b, Bro89a, Swe89a]. Solomon [Dre82b, Dro23e]. Solution [Abb79, Ano92o, Ano94h, Dow80d, Gla75c, Mac06, Sen47, Mul66, Sib73, Stu73, Ung78b, Agn82a, Ano90-34, Ano93x, Ano15e, BR15, HK96, KJE16, Mor87a, Per88a, Rot94h, SM90]. Solutions [Hol74, HWH78, Hur78b, McC09, Pom65, Tel73, Ano11e, Bac80, Huq11, Kap81a, Kre11, LBCA23, Myn11, Nye82, Pan23, Pif18, Squ21a]. Solve [You75, Dah77, Gib91a, Hes78, Kre85b, Tud21]. solved [vH79]. Solving [Stu69a, SG62a, La 88b, Kar80]. Solzhenitsyn [Red69]. Somatic [Bru58, Neu58a, Pau70]. Some [Ano49-30, Ano55x, Ark01d, Bas80, Bec64, Bor52b, Bru52, Cle48b, Cha52, Cha75a, Cha53, Com74, Cot52, Cou52a, Dav52, Day76h, Due21, Erd52, Fed59c, Fel84b, Fer52, Fer61, Fin77b, Fla65, Fri52, Gom61a, Gom61b, Gom75c, Gro70, Gug52, Had52, Hoa59, Hol75, Joh65b, Kal53, KD59, KM90, Kas52, Kat61b, Kel98b, Kuz58, Ler52, Lib59b, Lin12, Mal54, Mar65h, Maz69, McN59, Mel53b, Min52, Mon52, Mul73b, Nol90, Oel20, Oli52, Pei52, Pol52, Pri52, Rev62, Rod67, Ros60, RSW55, Roz64, Sci55b, Sad52, Sch52, Shi46b, Sim81, Ste53, Ste77c, Sto78b, Str05e, Tay47, Ter58, TR80, Wei59a, Wio77, Wri99a, Wya52, You66a, Bas86, Day77d, Dro22e, Eds76, Hec22, Hol77, Isa92m, NA90k, Tur77, Ano65e, Sto58]. Somebody [Fri74b, Leo19]. someday [HS13]. Somerville [Cra83b]. Something [Rot04f, Fie21b, Hof88, Pat80b, Woo92]. Sometimes [She97]. somewhere [Ano90-33, Pat78b]. Son [NA87i, Ark94b, Gra81a, GK94]. Song [Wai60, ER90]. Sonia [Bre93]. Sonic [Ano70d, DeN65]. Sonya [Mos93b]. Soon [Ano46u, Lau60, Mar02b, Gho11, Gor96d, Mar02a, NA90k]. Sooner [STS^{+}49]. sooth [Bas80]. Soothing [Kli59]. soothsayers [Bas80]. Sophisticated [Lev64]. Sopot [Rab67h]. Sørensen [Pat81a]. sorrows [ZK94]. Sorry [BR57, Fel70b, Kra99, Pan00, Rab59p, Rab61m, Fre93b]. Sort [Ano95l, Fel92, Sch04d, Ano95k, NA92a]. Sorting [Bul08q]. Sought [Has68]. soul [Dro17]. sound [Bar87, Rot99c]. sounded [Dro24e]. Sounders [Yor75b]. Soundly [Thi55]. Source [Ano48p, Men48, Pen77a, Sch47, Wal06, rWeKHF47, Gay98, JH16]. Sources [Alf74, Ano78k, Ano87n, Rap65b, Sch95b, BV20, Moo14, Mer79, War49]. Sourwine [Kal61]. South [Das96, Dje76, Akh17, Alb89a, AZ91, Alb93b, Alb93f, AH93b, Alb94f, Ano86q, Ano10a, Arn96b, AH22, BG14, Car17, Cha11, Con86, Flo88, Fra65, Fry85, GN10, Gor96b, Hea83, Hol85a, KHB^{+}05, Kan11, KKL15, Kim95, Kim20, Kna54, KR12, KN17a, Lan70, LL16, Lop76, Mag85, Man80, Mat82a, MR77, Mit83, Mti96b, Myr69, Nai81, NF83, NA87e, Pab88, Per95, Raj14, San91, Sar17, Sid00, Sig10, SJ68, Soa96, SS88, Sut04, TA94, Van22, Zai06, Zam95d, Dow80c]. Southard [Sto16]. Southeast [Lal67a, Van22, Woo05, Cal63, Kim22, Wei78c, Wei78d]. Souther [NA88a]. Southern [Pri81b]. souvenirs [Ano94o]. Sovereignty [Ber46, Bra47, Ste76a]. Soviet [Ack85, Ack88a, Ano77m, Ano86w, Ano87r, Ano88d, Ano90o, Art71, Bal88, Bei84, Bla85a, Bla87a, Bus85, Cra82c, Cra83b, Cra83d, Cra84b, Cra84d, Cra84e, Cra84c, Dar91, Ehr81, Eri95, Eva86, Fel73a, Fis83, Gal86, Gus83, Jon87, Jor87, Kin86, Klo85, Kol63, Leb94, Lon85c, Mac62, Men92b, Moo94a, Ora65, Rab72a, Ran85, Ric52, Row88, Ses83, Som91, Soy89, Uhl95a, Val86, Vuc69, Win90, You66c, Ack88b, Ame86, Ano47w, Ano48-40, Ano49-38, Ano50q, Ano52f, Ano53o, Ano53-41, Ano54-29, Ano54-30, Ano56c, Ano57k, Ano58j, Ano60b, Ano60j, Ano60r, Ano63z, Ano64a, Ano64j, Ano68f, Ano80m, Ano82c, Ano84i, Ano85y, Ano86-31, Ano87i, Ano87q, Ano87-43, Ano88s, Ano90-46, Ano91-59, Ash50, Bal82, Bei82, Bla85b, Bla69b, Bor87, Bro85f, Bru82]. Soviet [Cas85a, Cle82, Coc81a, C^{+}89, CN90, CN91, Cof70, Col88, Coo90, Cra82a, Cra82b, Cra83b, Cra83c, Cra84a, Cra85, Cra86a, Cra82d, Dal92b, Dal83, Dav57c, Dav51, DeW60, DK80, Duf86, Dun85, Ein48c, Ekl80, Elb81, Eme71a, EG86, Eri88a, Eri88b, Eva87a, Fau87, FUvH85, Fel77o, Fel79a, Fis83, FP90a, Fow58, Fra76a, Fre63, Gai69, Gar88b, Gar91, Gil68, Git54, Glu50, GB88, Gol82c, Gol84b, GR67, Gom65, Gor85, Gra82a, Gra83, Gra86, Gro47c, Gro48, Hai86, Han91, Han87b, Har69b, Hec86, Hek60, HK02, Hoc57, Hof86, Hol86c, Hol87b, Hol93, Hol94b, Hou82, Hou85, Hsi65, Isa90s, Jac74b, Joh60b, Joh64a, Joh64b, Joh83b, Jon59, Jos87, Kal85c, Kap85, Kap92, Kap78b, Kap78c, Kat90b, Ken78, Kla84, Kli52, KK85, Kok88b, Kor77, Kra57, KP85]. Soviet [Kre87b, Kre88f, Lal67f, LD89, Lap59c, LT99a, Lav55, Led85, Leg87, Lei89b, Lic77, Llu00a, Loc91a, Lon52a, Lon57, LP85a, Lov63, Luc82, Lyn88, Mac65a, Mar82a, Mar79c, Mar89c, Mar58d, Mar63, Mat85, Mat79, Mcc87, MS77a, Mit62, Mot89, Nat86a, Nat85b, NA87g, NA87h, NA87l, NA88c, NA88d, NA88f, NA88s, NA88r, NA88q, NA88p, NA89e, NA89f, NA89n, NA89p, NA89r, NA90b, NA90j, NA90m, NA90o, NA90n, NA90p, NA90q, NA91d, NA91g, NA91n, NA92c, NA92d, NA93o, NA94a, NA98e, NK09c, Obu89, Osn77, Par87, POF49, Pow82, Pru85, Rab47j, Rab47k, Rab49e, Rab52e, Rab52f, Rab53d, Rab57c, Rab61h, Rab62a, Rau92, RE68, Ric86b, S.46, Sch54c, Sch88b, Sem53, Sha86, Sha84e, She64a, She85a, Shi47a, Shi49, Shu87, Sig87a, SF81b]. Soviet [Ste85c, SH89b, Stu87, STS^{+}49, The90, Tim87, Tro58, Tsy83, Tuc66, Tur78, VFIS48, Vid90, VA11, Wei83a, Wei91a, Wei86c, WW80, Wil80d, Yaz88, Yem87, Yev86, You85a, Zag91b, dJ57b, vL86, Ano78e, Ano78n, Ano88c, Asp79, Cra83a, Cra84b, Ekl80, Ham61a, Ham61b, Kat57b, Lap59b, Rab65c, Stu89]. Soviet/Russian [NA94a, NK09c]. Sovietica [Fis90]. Soviets [Par88b, Ack87f, AHK90, BS85, Ded65, Dor85b, Gya90, Han90a, Has88, Jac79, KFT61, Lei89a, McK89, She85b, Uhl00, vH90a]. Sowing [Rab98]. soybeans [Cat87]. Space [Ano56d, Ano58j, Ano60m, Ano61n, Ano61q, Ano61s, Ano61t, Ano62-29, Ano63y, Ano67e, Ano67h, Ano68k, Ano86-36, Ano96l, Ano10a, Bal79c, Ber58, BB10, Bla67, Blo63, Bor58c, Bor66, Bou87, Bre59, JFB19, Che70, Cle67, Cor69, Cow65, Dad67, Deu83b, DuB63, Dys69b, Dys69c, Fed58, Foe07, Foz61j, Foz62h, Fri64, Fru63, Gal65, Gar81, GP84, Gol66, Gor67a, HWG95, Has67, Hav69, Hec06, Hol86b, HM72, Jac73e, Joh22, Kan71a, Kat58, Kre05, KP71, Lac85, Lal66c, Lew67a, Lew67c, Lew69c, Lig84, Lov63, Lov68, Mac67, Mar65d, Mar72b, Mec63, MGZH07, Mil63a, Mol03, Moo01b, Moo01c, Moo01d, Moo03b, New58, New61, OGP84, OR70, Pre58, Pit61, Rab61i, Rab63d, Rab66a, Rau69, Rub63, Sag73, Sch62a, Sch62c, Sch63e, Sim60a, Sim61, Sin70, Smi72a]. Space [Spi63, Spo64, Ste72b, Stu69b, Sut61, Ure67, Ure69b, Van61, Vel84, Woo01c, Woo61, Wri61, vdH61, Aft86, All16, Ano90u, Ano90-44, Ano93-29, Arb19, Arb85, Bro87c, BHK^{+}21, BB22, Fel84c, FP90a, GSB^{+}83, GN10, Gho15b, Hec19, Kor19, Lay79, Mar94, Mat85, McD85, Mec22a, Pal77a, Por19, RWK^{+}83, RH14, Rem15, Rob18, Rot04h, Sam22, SW21, Sok22, Tar93, Woo12, Ano63y, Ano86i, Deu83b, Dow81f, Fry65, Lac85, Lam69, Lew68b, Mic69, O’L73, Tho80d, Deu83a, Fri64, Mos90, Sch63c, Tho80d, Ano79a, Tir88]. space-based [Rob18, Woo12]. Spacefaring [Ano69n]. spaceport [Les93c]. Spaceship [Har72a, Lew72a]. Spacewar [Deu83b]. Spain [Ano91u]. Sparkling [NA90d]. Sparring [Wes98, Yaf06]. Spartans [NA89j]. späten [For80]. Spatial [Joh93]. spawned [E.77]. Speak [Ano60n, Ano60p, Ano65g, Ano66c, Ano66r, Ano90o, BAA^{+}75, Bur62b, Fed68b, Fei90, Hol93, Isa03b, Jos81, OW65, Sch54d, Tam71a, Ano12a, Dro23b, Fel80i, HRS81, NB78, Pai85d, Sto55c]. Speaker [Ano45c]. Speaking [Cro07, Mat98b, Ano96o, HC87, Lev80]. Speaks [All48, Ano55q, Bél48, Eis57, Isa02f, Khr57, Mor83b, OS00, Pik48, Tru48b, Isa90z, NA89a, Tru83, RI83]. Spearhead [Wei68c]. Special [AJSB79, Alb81, Ano70e, Ano71f, Ano72h, Ano78m, Ano80q, Ano95k, BEGJ72, Ber72d, Bey12b, Bin79, De 70b, Fel80l, Gan72, Gar81, Gen72, Gol78a, GPR81, HAIT80, Hor77, Hud78, IS67, Jac72a, Jac72b, Jac73b, Jac73c, Jac74d, La 79, LR79, Lew71a, LC78, Mil81, NA90n, Pug78, Pag68, Pan81, Rit48, Rob75b, RR68, Sac72b, Sha78a, Sha78b, Ano83n, Ano98o, Ark96g, Cra82d, DM24a, Man15, Mat82b, Cra82c, Cra83d, Fel78h, GM54, Tsi83h, Tsi83g, Tsi84b]. special-weapons [Ark96g]. specialists [Ekl80]. specialization [Hol84b]. Specialized [Ano59h, Ano60s, Ano61u]. Species [Fra72, Lor70, Mei51d, Ryb71, Wel60, Ano85i, Fei81a, Har81]. Specific [Lee73, You66a]. Specifications [Kan77b]. Specious [Bre82]. specter [Wat77]. Spector [Ano91s, Don86b]. Spectrum [SD10a, Buc83]. Speculation [Dys63a]. Speculator [BB63]. Speech [Ano46-77, Ano52j, Chu54b, Chu54c, FDC52, Fox52, Haf54, McM48a, Rab47j, Rab54h, Ano47f, Kat54b, Leb79, Bev56]. Speeches [Ano59e, HUC48]. speechwriter [Isa89d]. Speed [Cal65, Civ02, Dal03, Jac74b, Pri77, Ekl77, Ial18]. speeding [Zha20]. Speier [Bar58a]. spelling [Ano90-33]. Spencer [Dow82g, Smi88, Wil69c, Bul08o]. Spending [AW85, Isa97c, Isa99b, Isa00a, Isa04c, AG86, Ano86p, Ano86-29, Gar84b, Isa92g, Isa95a, Isa95b, Isa95c, Isa96e, Isa98b, Kau90, Mar89a, Mec21d, Ode88, Sea85b, Smi21, SD87, Ano78o, Sch80e, Siv80]. Spent [Alv02, Bas86, BWA85, Kan11, KKL15]. Spent-fuel [Bas86]. Speth [Ano11c]. Spezialinventar [For80]. Spiders [NA89i, NA90p]. spiderweb [GU88]. Spiegel [McC54c]. Spiers [Bur90c]. Spies [Ano51b, Bam00, Kum99, Mos90, Ric96, Ano92x, Kre89b, Pra92, Mos93b]. Spigelman [Her82b]. Spin [Ano96n, Smi89]. Spinning [Ano96m, Cha06b, Isa97f]. spins [Kre90]. spiral [Mor83c, Tsi83c]. Spirit [Bix49, Bri49, Key57, Lil49c, Mar49b, Sin56, Sta49, Was66, Ano95k, Dow91]. Spiritual [Cus48]. Spiro [Bal76]. splinters [TM96b]. split [Gam94]. Splitism [Moo99i]. Splitting [Lev12]. spoil [Rot92v]. spoils [Bro86e, Isa91g, Uhl98a]. Spokesman [Ame68]. sponsor [Ano87-32]. Sponsored [KND10, Lal65d]. spontaneous [Fre77b]. Spooks [Sie06f]. Sporn [Spi72a]. Sport [Abe74]. sports [Rot92v]. SPOT [Ric89, Rot93q, Zim87, Zim89a]. Spotlight [Ack88i, AH92l, NA93m]. Spots [Ano07h, Ano89j, Zim89c]. spotted [NA90p]. spotting [AH93a, Zim89c].

Spratt [Wee95]. Spread [Ano91s, Ano91-66, Bob65, Gol95d, GM09b, Hal75b, Ham93, Ing67, Kal64, Mil85a, SS90b, You66a, Ano86g, Str67]. Spreading [EB78, Mat98f, Sha98c]. spree [ZK94]. Spring [Cum92b, Dia63, Foz62e, Mar65a, NA91i, Ehr79]. Springer [Rot97c]. Springer-Verlag [Rot97c]. springs [Ano95j]. sprung [Rot98g]. spurned [She86b]. Spurr [Whi89]. Spurs [Moo91g]. Sputnik [Dic01, Dic11, Dic03, Ger58, McD85, She69, vH58, Gol78a, Moo02d]. Spy [Ano92w, Ano98s, Bou98b, Fly02c, Fly03, Gre55a, Kee56, Mos91, Rab51a, Rob88, Sch00e, Ure56, Ano97n, Flo90, Wei89b]. Spying [Ark88b, Zim89a]. Spymaster [Eld92]. Spyros [Lyn78]. Spyship [Ark88a]. SQI [Gas93b]. Squad [BF04]. Square [Ano80o, Cle04b]. Squaring [Bha99, Led71b]. Squeeze [Nad98a]. Squeezing [Dav93]. squirms [Kom90]. SRAM [NA87i, NA90l, NA90p]. SST [MH67, PvH72, Shu72b, Shu72a, Smi74a]. St [Ano84d]. Stab [Nus80]. stability [Ano98o, Asp87, Ing85, Mar85b, Nit89, Vad21, Vel84, Yua23, Gle94]. Stabilization [Bre69a, Gal78a]. stabilize [Zha19]. Stable [Ano48-38, Cap77]. Stacked [Smi94a]. Staff [Ano46-94, Ano74j]. Staffs [Got64]. Stage [LW69, Mil52, Szi60a, Car03]. Stages [Ano69s]. Stagg [Moo92l]. Stähl [Kat52b]. Stakes [Kin86, Kre87a]. Stalemate [Cav55, FGJ65, Her85, Hym71, Kno55b, Lev62, Mei56, Szi60a, Fel77i]. Stalin [Gol93d, Ano47w, Ano90-61, CNB95, Mar48e, Szi47d, dW90a]. Stalinists [Zag95a]. Stalking [Lew73a]. Stalled [Dei05]. Stalling [Pri76a]. stalls [Gol79]. Stamp [Mun54]. stamps [Ano91-41]. stance [Eps87a]. Stand [Ano50r, CO06, Eps94, Hor71, Pra03c, BW18, Cle87a, Isa89f, Lor77, Pat80b]. Standard [Ash76, Wal87b]. Standards [Ano63x, AM74, Bar63, Ham72, Rot79a, vL48b, Alv84, Ano77l, Nus85, Ano53t, Ano53x, Sac46]. Standing [Sha87a, Kel10]. Standoff [Per95]. Stands [Str07e, Ter04, Wil69a, Fel82d]. Stanford [Bou02, Gre82b, Ano11h]. Stanley [All55l, Ber78a, Her90b, Sil79a, dF79d, Tsi84a, You66b]. Stansfield [Eld92, Moo97b]. Star [Ano86i, Cla91, Est01, Han92b, Har70, Hit07, Lac85, Mec19h, Ric86b, Sar85, Smi87b, Jon91a, Jon93a, Sch81c, ATL85, Aft86, AK98b, Ano97k, Arb85, Ark86g, Bal86a, Bro85i, Bro87a, BS85, Fel85c, Gla85, Gom87, GK94, Gur88a, Gur88c, Har86d, Hec19, Her87, Isa88c, Isa93g, Isa95d, Isa96f, Jon93a, Ken85, KW86, Kub88, Lan90a, Lea87, Led85, Men96b, Mon89, Mye86, NA87f, NA89j, RK87b, Rot92o, Rot94k, She85b, She86a, Smi87b, Smi87c, Smi00, Swe87a, Swe87b, Tho80d, Tir86, Win87, You87b, vH85, Gol88g, Lee91a, Pre86]. Stares [Tir88]. Starke [Roc85]. Starr [Mye89]. Starry [Ano90-31]. Stars [Bri89, Rab62k, JFS20, Rot98e, Jon91a, Jon93a, Dra69]. START [RKK22, Bre69a, War88, Ber23, Fal79, Fra79, Kis79b, Nov77b, Opp90b, Rev81, RTB^{+}19, Sal79a, Ade91, Ano91-68, Arb91a, Arb91b, Arb93, Coc89, Dow82i, Dre91, For91, Ham90, Han96a, Isa90t, Isa91p, Isa96c, Kor90, Kra91, Kre83b, LP91, Lin84b, Loc91b, Loc92a, Loc92b, Mar91b, Men89, Men91b, Men91c, NA88o, NA88k, NA90l, NA92j, Pai82e, RD91, Sig88b, Tel91, Tsi88, Wal89c, War91]. START-II [Arb93]. Starting [Wei91a, Kre88f]. Startup [Nel93b]. Starving [Wor58]. Stash [Mae01]. Stasi [Ano99a]. Stason [dK60, Lap57]. Stassen [Ano47w, Ing55b]. State [All50, Ano60f, BN67, Bra74, Day97, Dob52, Dre51, Dub55a, Gui07, KND10, Kli52, Klo85, Lef77b, Lei72b, Lon52c, Lue96, Mac84b, Mar48c, Mil69b, Nie66a, Rei63, Rub08, Shu78, Sin78, Ske90, Sly99, Ste65a, Ano97j, Doy53, Dro22a, Dro22c, Eri95, HIKWB10, Moo93i, Sha22, Ust93e, Woo97a, Adv50, Ali57, ABB^{+}46, Ano46-54, Ano46-81, Ano46-89, Ano46-88, Ano46-95, Ano58d, AC65, CCMU46, Eli46, Hor66, Ing57a, Joh67b, Kil65, Mor55, Roc85, Shi46b, Tel46b, Tod64b, Ure46c, Eis67, Fel77o, War86, Roc85]. State-Sponsored [KND10]. statehouses [Fie24]. Statemanship [Hof61]. Statement [Ato54a, Ano46-89, Ano52y, Ano53n, Ano53-42, Ano55i, Ano57i, Ano57k, Ano57j, Ano58g, Ano58l, Ano59j, Ano60b, Ano61i, Ano61k, Ano62u, Ano62x, Ano67d, Ano67l, Ano67n, Ano71g, Ano72h, AIL^{+}62, Cou48, Cou57, Che52, CH54b, Con55, Day74a, Dea53b, Eme48, Eis58, Fed62, Fed64, For49, Fra65, Gro49, HKL^{+}49, Ken71, Lif69, Mul67, OBC^{+}49, Pug83, Pai82j, Rab61j, Rad53, Rus55c, Sal53, Cou50, Str54b, Tru50, Tru51, WG69, Wol50, Zim48, Ano64p, Ano85f, Ano85g, Ano85i, Ano96g, Bro23, Hol88, Ano65g]. Statements [Ano47w, Ano48s, Ano67o, OJCP^{+}47, Sch77a]. States [Abb85, Ano47z, Ano48-28, Ano48-29, Ano52b, Ano57d, Aus47, CB95, Dew55, Eba61, Fel83e, Gol69b, Hut51, Kor92, Lew61b, Mer90, Mes84b, Moo95b, Mos84, Myr69, Nie63, Rei63, Roc81, Rya07, Sca91, Shi61, Sla88, Sol68, Ste80a, Uhl95b, Van88, Whi51b, Wio84, Woo56, Wur68, ES89, Kri87, McD13, AK98a, Ano54-27, Ano60e, Ano60d, Bro47f, Brz12, Cot19, Dal22, DRM^{+}09, Dje86, Dow82e, Ekl80, Eme71b, Gom83, Gru70a, Hol84c, Hum58, Hun85b, KL65, Kar17, Ken83, Kno60, KK19f, KK20d, KN16d, KN17d, KN18d, KK21f, KK22e, KK22f, KKJK24e, Lap51a, Mau50, May17, Mec22c, Mec22b, Mia16, Mor50a, Mor54, Mor67, Nit57, Rav83, Ros82, RE19, Sen57, Seg13, Sha87a, Sha78b, Shi57, Sig10, SL83, Smi21, SSR83, Szi47a, TvH17, Tha00, Tra72c, Wei53b]. States [Wei78c, Wie85, Wu23, Yua23, Zha19, vH16, Dow81d, Ekl80, Gre58, Nad79, Per95, Smi54c, Dow82a]. Statesman [Cor53, Sch91]. Statesmanship [Com63b]. Statesmen [Bor52a, Dea54b]. Station [Abb91, Spi60, Ste91a, Klu22, Ano57b]. Stations [Fri54, Rau69]. Statistical [Siv75, Zei73, AK81, RS16]. Statistics [Bro74, Com75c, Ney73, Ric52, Wei69b, Mei90]. Stats [Bul02b]. Status [Ano46-36, Ano46-56, Ano46-75, Ano46-96, Law69, McL11, Mul66, Pac59, Rab59d, Sch08b, STFK09, CM12, Man85, Mar76, Orl84, SFT11, SF12, SF14]. Stay [Cum01b, ME51, Par78a]. Staying [Hol01, Sha96]. Steady [Bra74, Alb93a, Ano91-34]. Steal [Sto99, Lum94a]. Stealth [Ark89d, Bul07d, Ano90x]. Stealthy [NA87e, Spa95e]. Stearns [Com48b, The48]. Steel [Rab69b]. Steelman [Stu48]. Steen [Met83]. Steer [Mul68]. Steering [Ano96-27, Bet75b]. Stefan [Lew62a]. Stefanick [Ano87q]. Stein [Kra94, Sar85]. Steinberger [Moo93b]. Steinbruner [For87]. Steiner [Nor64, Mul73b]. Stello [Lan90b]. Stenehjem [Gra93].

Stent [dF79g]. Step [Ano63l, CLR75, Fel84d, Lan90g, STC74, Szi55b, Ano85j, Ano13b, Hen96, Isa89c, Kea24, Mal79, Men92a, Oel12, Sto97, TM94b, Wur93d]. step-by-step [Kea24]. Stephane [Smi68]. Stephen [Ack85, Ack97, Ano87l, Ano90o, Bal88, Boy91, Bus85, Cra83b, Das96, Moo96b, Par88b, Rei78, Roc81, Smi90a, Tho80g, Wei85b, Bul09c]. Stephens [Sti83]. Stephenson [All54d]. Steps [Ber81, Car76, Dam62, DS12, Fed58, Isa01b, Lan90g, Lon64, Sco72, Ste58, You63a, Ben76c, Eps96a, Fel78c, Hib95d, Kel84, Lin12, McC88b, Rot94a]. Sterile [dL68]. Sterling [Saw81b]. Stern [Dow83a]. Sternberg [She61]. Sternglass [Ano70b, Dys69d, HI72, Sag69, Wei69b]. Steve [And83, Ano90n, Dow80j]. Steven [Cha79a, Dow82g, Ede89, Hun87, Kre90, Kri85, Par86, Dro16a, Dro23a]. Stevenson [Rab65b]. Stever [Jas81]. stew [WS95]. Stewardship [Gar97c, Gra94, Kat95, ZM96]. Stewart [Diw79, Lev73, O’T00]. Steyer [Dro21d]. Stich [Tho80g]. stick [Fei90, Pat90b]. Sticks [Gro07]. Sticky [NA91g]. Stiff [Men95, SV86]. Stiff-arming [Men95]. Stiffed [Veg97]. Still [AK04, Ano90-34, Ano97j, Ark97e, BPTB76b, BK81, Goo03, Kra10, Men99, Men00, NA93p, NA93t, Sie04d, Wur94c, Ano91y, Ark94e, Ark95d, Arn97a, BMPd89, Che82, Cot19, Fel82d, Flo88, Gar84a, Gev93, Hus24, Isa91b, Mat98a, Nat81, Nor94b, Sch76a, Smi16, Ste89f, WLP97, vV23]. Stillborn [Woo00b, Woo00g]. Stilled [Ano73d, BBH^{+}73, Fel73d, Lap73b, Mei73]. Stillman [Sha54b]. Stimson [Ano96-27, Ber93b]. Stimulation [Jac72a]. Stine [Deu83a]. sting [Hib96a, Szi65]. Stinger [Ano92x]. Stingers [Wit03b]. Stiskin [vH77a]. Stöber [PF75]. Stock [Hux51, Ack88j, Dro16a]. Stockholm [AGC^{+}88, Ano72h, Bar77c, Bar78b, Bar79a, Bar80b, Bar80e, Bar80f, Bar80g, Bar80h, Bar80i, Bar81a, Bar81c, Bar81d, Bar81e, Bar81f, Bar81h, Bar81g, Bar85b, BEGJ72, Ber72d, Bor85, Bor86a, Bor86b, Ehr72, Gan72, Gen72, Goo85b, Jac72b, Jac73f]. Stocking [Wie87]. Stockpile [DM84, Gra94, Lar84, McM49b, Mil10, NA89e, NA93d, NA93q, NA96e, NA96b, NA97a, NA97d, NA98f, NK07f, Ros82, Smy51, Ano87-45, NCA85, NA87a, NA88d, NA88t, NA89q, NA89p, NA89r, NA90b, NA90r, NA91d, NA91p, NA92d, NA92p, Nor94a, NA94c, NA94g, NA95b, NA95e, NK13]. Stockpiles [AK04, Myr75, NA00c, NK02, NK06a, NA93c, NA94a, NA97b]. Stockpiling [NAD^{+}47]. Stocks [GM09a]. Stockwell [Wyl56]. Stokes [Dow80f]. Stokley [Fer57b]. Stole [Gre64a]. Stolen [Arn22, Lum94b]. Stoll [Mye89]. Stolper [Kon67]. Stone [You66a, Ark97d]. stones [Moo92c]. Stonewalled [Pra03b]. Stonewalling [Sch77a]. Stood [vH01]. Stop [Ano85q, Boy05, Con58, Dre87, Har01, Ing59b, Owe87, Rab58e, Rab64e, Rau96, Szi60c, Ano91-29, Epp20, Fel77j, Gei16a, Hop77, Pri89, RK93, Shi47d, Smi85a, Tsi88, Vil75, YWS20, ZA04]. Stopgap [Kel62]. stopped [Eve88a, Moo25, Rya89]. Stopping [Ano90-54, Fra90a, Gub14, Hog86, Ing57b, Isa88e, LST10, Rab73g, Mat97d]. Storage [JvH09, Ste73a, Wal84, Bas86, DL21, Joh80a, Kan11, KKL15, RG13, RL14, Twi23]. Store [BES02, Kem97, Wei97]. stored [LP91]. stores [Han98a]. Stories [Ano52i, Bri86, Mas99a, Mec19g, Rot02d, Ano92y, Ano94u, Diz94, Huq11, Kre11, Moo94j, Myn11, Swa86]. Storing [BWA85]. Stork [Tyn75, Ano96j]. Storm [Bar79d, Bon53a, Jac15]. Stormed [NA92a]. Stormen [Bar79d]. Story [Adl69, Ano51u, Ano53u, Ano57b, Ano78k, Coh69, DN67, For51a, For51b, Han92b, Han01, Leb92, Lei92, LM71, Lib48, Mar54a, Mor47a, Nie66b, Nor02a, Nor04, Rab52d, Ram88b, Rot05a, Smi68, Sto66b, Zei73, Abr86a, Abr86b, Ame86, Ano92r, Ano13c, Ber86a, Bru15, DH86, Fis86b, Hof86, Man86, Mar83c, Mat97c, Mat98e, McC86b, McC86a, Mil10, Osa77, Pat86c, Pat86b, Tan88, Tho86a, TB23, Tis90, Wei86b, Wei86a, WM78, vHC86, All59l, Ano97k, Bro47c]. storyland [Ark95e]. Stowe [Ken61b, Khr61, Rab61k, YTS61]. Straight [Sch05d, Ack88d]. Strait [Yua23]. stranded [Lei82a]. strange [Isa98a]. Strangelove [Goo04a, Goo04b, Ano05b]. Strangely [Ano52z]. Strategic [Ano46f, Ano48-32, Ano49-40, Ano86k, Ano96v, Ano96w, Ano98r, Ano99-27, Bar73a, Bre69a, Bro57c, CO85, Dot65, EB73a, Fre84a, Hob85, Joh79, Kap82, La 78, Lal65g, Lon68, Mil68, Mor62d, Mye87a, NA90s, NA92l, NA96f, NA96d, NA97e, NA98g, NA98d, NA99c, NA99d, Nor02a, OSS84, Orn85, Pan71, Pol09a, Pos85, Roc81, Rod69, RKK22, Sar85, Sco72, She83a, YG77b, dF79h, Ano80i, Ano87-43, Ano92-31, Ano99-29, DG07, Fel80k, GJ84, Kal85b, Kin78b, LP10, Mil10, Mit97, NAC87, NA88s, NA88r, NA89n, NA90o, NA91n, NA92c, NA92i, NA93n, NA93r, NA94d, NA94h, NA95d, NA95f, Pai80a, PB^{+}01, Stu87, Tsi84b, Tur98, Uni78, Vel84, Ano51q, Bai92, Roc81, Roc83a, Ano84h, Ano87m, Ano87q, AM86b, Gar72, Her85, Sar78, You66c]. Strategies [Abb84a, Amd65, Bar74, Lov74b, Lov74c, Per87b, Sch78a, Hen17, Mil12b, TS22, She83b]. Strategies/Part [Lov74c]. Strategists [JL74, Pil82]. Strategy [Ano64k, Ano85c, Bar88, Bet62b, Bre69b, Bro62a, Cha69c, Cle67, Cop65, Cra83b, Dow82b, Dow84d, Eng08, Gal56, Hul98, Ing56a, Kal63a, Lap50, Lap63, Lar62b, Leg56, LM61b, Mar89e, MF63, Mic64, Mil63b, Mor54, Nel84, Rab56l, S.55b, Sch60b, Sch66, Sch80e, She61, Siv80, Ste86c, Sto62, Tam71b, Tho80b, Wal76, You65a, You66a, Ano84l, Ano87i, Ano87-33, Ark83d, Bar87, BT82, Bro85g, Bro85h, Bro86f, Bro86g, Bro86h, CS17, EE81a, Fel82f, Hat85, Hol77, Kin78b, KE17, LaF90a, Man15, Mec19f, NA88f, Osg80, Pol79b, Pos87, Ros87b, RW80, Tal18, BGR61, How65a, Jac80b, She60, Ack87a, Ano85b, Ano87q, Deu83b, Gan84, Har85b, Mas66, Roc83a, Tra90, Bre65, Nic54a, She60, Sta64, Wol63]. Strauss [Mul66, Bec78, Str54b]. Straussman [dF79d]. Straw [Ner76, Zag95b]. strawmen [Gor95e]. Street [Wil76a, Pat81b, Joh01]. Streetcar [Lew68a]. Streets [DeB81b]. Strength [Ber53b, Kin86, Mar55b, McN64, Ano88s, Cra83b, Hil65b]. strengthen [Fet87, Sen87, Sha10]. strengthened [Nov77b]. Strengthening [BB20, Isa93p, KND10, Len11, Rin66, Ber84a]. strengthens [Dea95]. stressed [Ing76, Van90a]. Stresses [Gor51a]. stretch [Dal92a]. Strickland [Sch69d]. Strike [HC83, Lap72, Mem79, Mat98e, NA92k, Pol09b, Ram88b, Ure48c, Zam96c, And81, Ark85h, Ken79, KK98a, NA87d, Pei84, Ray90, Tho85a, Dow83d, You84b]. Strikes [Nan73, Ark96d, Cal17]. striking [Mar89c]. Strobe [Cra84e, Sto85]. Strom [Isa95c]. Strong [Jac73f]. stronger [Dea95, Sim91]. Strontium [Bri57, Lap56d, NW70, Ore62a, Ore62b]. Strontium-90 [NW70]. Structure [Ano46-97, Ben71b, Kow49, Opp59a, Str67, Kal77a]. Struggle [And71, Ano69q, H.84a, Val86, Wad74, Hud77, Jac92, Mia15a, Tre18]. Struggles [GB88, Uhl96, Ack88a]. STS [Ano83q]. Stuart [Ano78l, Jac80b, Smi75c]. Stubbing [Ada87]. stubborn [Fig23]. Stuber [Har79b]. Stuck [Mac06, Woo01c, Isa90q, Kis93d]. Student [Ano87n, Cah71, CN88, Gre67e, Har56, Lew71e, Lyn69, McG68, Rab68e, Sel69c, Ano15d, Gan83d, GH85, Kal82g, Lan90n]. Students [Abe74, Ano46r, Ano50y, Gil55, Iff67, Joh83a, KAT^{+}71, Pow59, Wol55, Zwe85, Ber23, LM80]. Studies [Ano56h, Ano69o, Ano84h, Bac58, BAB^{+}46, Jr16, Hol58, Kuz58, Mic69, Osb58, Phe56, Roc83a, Sch68d, SS73, Soc15, WKW57, Wyl57, GL90, Lin84d, Nus85, Per90, Was82, Ano46-69, Nei82a, Pan63]. Studnet [Ano79n]. Study [Ano55r, Ano56n, Ano63y, Ano70e, Ano81f, Ano97r, Bar66a, Ber84a, Bet75b, Car72, Day75b, Deu68, Duf87, For83, Fre84b, Git54, Hyn69, Kel79, Kol63, Len60, Lew70e, Man04, Mar94, Mil69c, Mun54, Mur70, Nor64, PWW63, Ras75, Sai72, Sch69a, Sch75b, Sch80e, Sei69, Sim57a, Siv80, Sut54, Tho82, Tod78b, Wyl56, Alp92, Ano90z, Bun82, Gre77, Gre23, Hol77, RS16, SL13, Sin79a]. Studying [Mel75, Sto19a]. Stuff [Hol05b, Sch07d, LL97, Lym97]. Stump [Kuz04]. Stumping [Ste07b]. Stung [Ano92x, Wit03b]. stupid [Ark96a, Isa93e, Sch13b]. Style [Voh71b, Ano90-35, Ano91-40, Vid90]. Styles [Mar89e, Sel69c, Fer88]. Stymied [Foz61k]. Styron [Dow82e]. Styx [Rot89a]. Sub [NA88h, BP90, Lei82a, Lei89b, Yud85]. Sub-Saharan [Yud85]. Subcommission [Noy47]. Subcommittee [FCC^{+}46, Ano55y, Soc55, Spi57b]. subcontinent [HK96, MN96]. Subcontinental [KP98]. subcritical [dTdLF24, vHJ96]. Subculture [KAT^{+}71, Ham76]. Subcultures [KAT^{+}71]. Subdued [Ols62b, Ols62c, Opp62b]. Subject [Mel75, Tan06]. Subjects [Alb57, Cla90]. Sublimate [Lor57]. Submarine [Ano88c, Gol88d, Lap54c, Ros75, Ano93z, Cla16, Han87a, NA90d, WF16, Mar94]. Submarines [Ano75d, Ano93-36, Cra67b, Oga68, RC68, Ano96j, Cot19, NA87b, Pea90, Rho90]. Submerged [Amr54]. submerges [NA88f]. Subramanyam [Ano95w]. Subrata [Ano10a]. Subs [Mil02, Tsi89b, Han93]. Subscription [Wit52a, Wit52b]. subsidiary [Ark87b]. Subsidies [Bra17, Rom77, You47]. subsidizes [Fer87]. Subsidizing [Gru78]. Subsidy [Mos67, Abb82]. Substance [Ken66]. Substances [Tho80b]. Substantial [Lal65g]. Substitute [Cal80, Fel79g]. Subversive [Rho90, vL65]. succeed [Tro20]. Success [Ark94c, Nat86b, Wal89a, War91, Bes94, Bri80a, Eps90b, Kow77, Nei88, Pan88, Dow80i]. Successes [Fis83, Isa90i]. Successful [Sha69, Ano53o, Sel21]. succession [Hou82]. Such [Bou02, Off02, Sch84a, Tsi83b]. Sudan [CO06]. Suddaby [Dow82f]. sudden [Uri94b]. Sudoplatov [Les94, McM94a]. Sue [Sca91]. sues [Veg97]. Suez [Fin65]. suffer [Abr86a]. Suffers [Gol95c]. Sufficiency [Sto68b, Aye85, KDFS81]. Sufficiently [McM48b]. Sugarcoating [Ahm99]. Suggested [Ano46s, Tel46c]. Suggestion [Ano46-30]. Suggestions [Gil77, Sen47, SS73, Wyl56, LB78b]. Suggests [Ano49q, Spi49, Wil62, RS16]. Suicidal [Fel81g, Lev77b]. Suicide [Ale60, Tho79a]. Suisse [G.84a]. Suit [Bro55a]. Suitcase [Chu05]. Sukhomajri [SJ82]. Sum [Goo05, Ric01a, Ano97k]. Sumit [Das96]. Summarizes [Sci48]. Summary [Ano46t, Ano47f, Ano51b, Ano53d, Ano53t, Ano55y, Ano56l, Ano66l, Ano71g, Ano73h, Par66a, Rab60h, Shu78, Sim47, Tur52, Bro80a]. Summer [Ano83r, Gus99, Lew69c, Rab56e, Ano80p, Isa82c, Kin81a, LG15, Sch76a, Ano81l]. Summerson [Rab60h]. Summing [dF79b, Koe79]. Summit [Ano97s, Khr10, Kib12, Nay12, Rab72e, Squ12, Ano85r, Bro86c, Bro86e, FC88, Isa90b, Kal88, Moo96m, Nat88, Pan88, Rot88, Sak88, Sig85a, Sig87c, Tho88, War88]. Summitry [Kre90, Isa86f]. summits [Bla85b, DPT17, Kib12, Tho88]. Sumner [Dow80h]. Sun [Ano78n, Gol61a, Han95a, Lew64b, Mes85a, Rho95, Ste57a, Ano95n, Moo92a, Fri74b, Ano48s]. Sung [HC12]. sunken [OP98]. sunny [Fri20, Moo96h]. Suns [Ham91c, Tho80d, Bet58]. Sunscreens [Isa90u]. Sunshine [Lap59h, AM21]. Supek [Ded60]. Super [Ano47n, Ano50-37, Pri86a, Thi50, Tob77, Ano98o, Cro86, Ano50-46, Mat98d]. Super-Paranoids [Tob77]. superbomb [Yor76b, Yor89, Zac76]. supercomputer [Mar97b]. Superfund [Gas93a]. Supergun [Har92a]. Superiority [Lal67g, Utl51, Bet81a, Bor57b]. Supermarket [Har86e]. Supernova [Bet81b]. Superorganic [Dob64]. Superphénix [Alb84, Ano84p, Ano97n, HA85]. superplane [Lan90e, NA89a]. Superpower [Ano87-39, Jam86, Loe86, LP85b, Mes83b, Mül85, Dub85, Leb85, Men97, Rub87, Jor87, Moo93a]. Superpowers [Bar85b, Eps88b, Jon91b, Kru74, Lei74, Kal85a, Kat90a, Myr80, Tsi76, Ano90p]. Supersafe [NA92e]. Supersonic [Cal66, Hoh66, Lun65]. Superstar [Tuc77]. superstore [Har93]. Superweapon [Bri89]. superweapons [Mes91]. Supplement [KS58, Rot58a]. Supplications [Hal76a]. Supplier [AH93a]. Supplier-spotting [AH93a]. suppliers [FKM87, Har87c]. Supplies [Nin59, Spu73, Zha15b]. Supply [Bat61, Cha68b, Spi72a, Abb12, Jas24, Lan89f, Lou90, Ram13b, Smi13]. Support [All48, Ano48-37, Ano78g, FP89, Jan80, Kay66, Kis66, Kli80, Rid47, Ros68b, Sub70, Ste46, TB23, Uni54, VFIS63, Wal51, Ano85i, Fie19, Ito14, Joh58, Lin86, Per20, Shu78]. supported [Wee95]. Supports [Ano46-47, Ano47p, Ano62-28, Wal46b]. Suppressing [Gut93e]. Suppression [AC76, Dob49]. Supra [Sch71b]. Supra-Nationalism [Sch71b]. Supremacy [Oli48, Jon87]. Sure [Pan22]. surely [Moo92a]. Surf [NA87f]. Surface [Tur73]. surfacing [Fie90]. Surfin [Ano95m, Mat97b]. Surgery [Arn90, Fav52]. Surgical [Con81b]. surpassed [Zag95c]. Surplus [Ano46r, Hol94a]. Surprise [Bak06, KHB^{+}05, Kre87f, Lal64g, Sch59a, KGS25, Sok91]. Surprises [Hel02]. Surprising [Kuh75, Sch75e]. surveillance [Ban22, CET77, Dro22a, Dro22c, JPR90, Ric86a, Sha22]. Survey [Alb73, Ano49n, Ano49-40, Bro74, Fis46, Lla70, Wyl56, Ano46f, Dro20a, Fuc70, Ber73c, Soc54]. surveyed [Zwe85]. Surveying [Wol01a, Wol01b]. Surveyor [Tur73]. Surveys [dRMT^{+}47]. survivability [SA82]. Survival [Adv60a, AC49, Ano71e, BW59, Ber71, Eme48, Fei81a, Fei71, Fow59, Hea83, Ing53, Kas77, Ksa91, Lew72a, McL53, McM52, Mol62b, Mor83a, Rab47g, Sha73, Shi66, Smi83, dF79d, Ben76c, Gei81, Joh81b, Pug88, Spe89c, Air55, Dow84b, Ano50-41, Fri72, Hul78, Sta64, Bla85a, Tho79a]. Survivalism [RS83]. Survivalist [Smi83]. Survive [Ack09, Jac74d, Ste69f, Szi60a, Szi63b, Dro22e, Fel79a, Fin92, Gol86a, Rab60g, Smi83]. Survived [Bru59]. Surviving [Bro63a, Got87, Lym11, LL83]. Survivor [Nus85, Osa77, Swa86]. Survivors [Aki05, Dow82e, Alv84, Ano15f, Blu23, Con81b, Pet11, SK84, Dow82g, Joh81a]. Susan [Ano91s, Bar90c, Lyn69, Dro23e, Sto16]. Susanna [Dow82e]. Susantha [Eas85]. Suspension [Exe54, McN48]. Suspicion [Kih47f, Ste07a, Osn77]. Suspicions [Ing59a, Agn77a]. sustainability [KH13, Smi13]. Sustainable [Ehr83, MR09, Teb85, Not15, RF21, Boo22]. Sustained [Rab05b, vD78]. Suter [McG84]. Sverdlovsk [Ano83a, Gre83b, Zil83a, Zil83b]. Swagger [Nas06]. Swain [Cog82]. swampland [Chu76]. Swan [Ano78n]. swap [Kib10]. swarming [Lac17, Sch18]. Swartz [Ano88x, Ano88y]. Swear [Hib01]. Sweat [Chu05]. Sweating [Sch07d]. Sweden [Int67, AJSB79, Ano62-30, Ano86o, GJS80, JS79a, Joh86, Lei89b, Lun59, Lun79, Mal63, Obe74, Pie16, RF21, RAB^{+}80]. Swedish [Ano88c, Tra71, Bar80l, Tho86b, Tis52]. Sweeping [ZF03]. Sweet [Elk62, Ing74, Lum94b, Don89]. swept [Zel76a]. swift [Isa92b]. Swim [Car77, Isa85f]. Swimming [CYP14]. swing [Gau24]. Swiss [Ano96o, BV21, Tsc87]. Switch [Fri74b, Jac74b, Moo05b, Ano91-41, Ano91p, Ano95i]. Switzerland [Llo71, Not15, PG79, Rot03f]. Sword [Asi63, Hir88, Isa92b, Wei80b]. Swords [Day77n, HHH07, Mar97a, NA90a, Wur97]. SYFS [Moo93j]. Symbol [Ken66]. Symbolic [Tay98]. Symbols [Tay62]. Symmetries [Fla68]. Symposia [EBC83]. Symposium [Ano53-43, Ano54-31, Ano71e, Dan72, Dje72a, Dob52, Fre72b, Gan52, Gla60, Hoa72, Kli52, Lon52c, Ney73, Rab56k, Rab56l, Rab56j, Ram72, Ric52, Sin67a, Fra61a, Dam62]. Symptom [Dou69]. Symptoms [Ein52b]. Synchronized [May52]. SynCom [Lew63c]. Syndrome [DeN76a, Car83a, Mol96, Sic96a, Dow81g, Lew63c, Dow79d]. Synthesis [Dan72]. Synthesizing [MZ07]. Synthetic [Lau76, McP65, SW10, OGFM18, SSK10, Sol80b]. System [Ano67i, Fra74, Gre55b, Hoo63, LLH^{+}62, Leg57, LEH^{+}71, Lew69b, Oom46a, Pit67, Rin57, Sie06a, Ste75b, Van90b, dF79d, vB69, Abr83, Ano87-38, Bud22, Eng78, GP18, Gre84a, Ham76, Hol82, PD12, Rob18, Whi79, YWS20, Gro56b, Rei78]. system-wide [Gre84a]. Systems [Cis52, Dot65, Ing68b, Lap61b, Mil68, Sch68d, Sha54b, Tal65, Ano78p, Bre13, Fin77a, Her79, HCH16, IES79, Jor18, MS77b, Mor84c, PN09, Shu76, Ano78n, Sha54b, Ano78d, Bro79b]. Szasz [Don86b, Mes85a]. Szilard [Fel81m, Gab73, Rab64b, And74b, And74a, Ano50n, Ano79f, Ano79g, Ano79i, Ano79m, Ano94q, Bes85, Gab73, Gol92b, Har65, Hut61, Lan92a, LS94, Mar48e, Szi47c, Zei70].

T [Ano75a, Ano78l, Ano87-32, Ano93p, Ano10b, BGR61, Boy95, Gab73, Gre82b, Har86e, Hec70, Her88b, Ing82, Lon82, McK74, Mos90, Pon67a, Rab72a, Rea80, Rot97b, Sco74, Sim85, Sto55a, Tho80d, Tsi72a, Wal85a, Ano90-35, Ols74b, Ark98a]. T-55 [Ano90-35]. tab [Isa93o]. Tabarin [Fuc70]. table [Ano78k, Ano78n, Ano78l, Ano78o, Ano78m, Ano78p, Obu89, S.80, Tho80g, dF79f, dF79h, dF79c, dF79e, dF79b, dF79a, dF79g, dF79d, dF80]. Tables [HH75]. Taboo [Tan06, Lew73a]. Taboos [Wil53d]. tabs [Yar80b, Yar80c]. Tachistoscopy [Bac59]. tacit [Bil85]. Tackle [Bul08a]. tackles [Pat80e]. Tactical [Cou52b, Gav51, Ing52b, KK19e, Lor92, RKR55, RK55, Daa87, Fit11, NK11, Sau10]. tactics [Boy86a, Cal86a, Dal17, Isa97e, Lan10, Cra83b]. Tad [Mos92]. Tag [Ano51-27]. Taguchi [Ano15f]. Tahiti [Cab97, Dan90c]. Tail [Alv87]. Tailings [Com75e]. tailored [Ban11]. tails [Rot92i]. Tailspin [Isa99b]. Tainter [McK25a]. Taiwan [AG98, Cot22, DE23, Dre99, Dro23g, Gol23, Spo69, Wu22, Wu23, Yua23]. Taiwanese [Kem09]. Tajikistan [Ano94w, Zag94a]. Takash [Amr51]. Take [Ano46-72, Cle99, Ded65, GRA^{+}76a, Lew69b, Smi90b, Tha00, vHJ96, Ano90-31, Eps97, Fel80c, Fel93, Fra79, Gol88f, Has88, Isa95b, Lor77, Uni78, ZC92, Ano93u]. taken [Ber85, Seg13]. takeoff [Smi92c]. Takes [Gol92c, Pal64a, Zim48, Ano94m, Bar90b, Gol90e, NA93s, Sch05a]. Takeuchi [Ano15f]. Taking [Ack88j, AH02, Ano95i, Ano13d, BT94, Dro16a, Gus07, Hux51, KW86, Kra73, LST10, Moo97q, Pit91, Rec89, Str07e, Wri04, vH89, Nil14, Ano94s, Nei82a]. Tal [Dow80i]. Talbott [Cra84e, Sto85]. Tale [Har88c, Rob78a, Van60, Woo02, Ark95b]. Tales [Ark95e, Gus01, Tel49, Hib95b]. Taliban [Wit01a, Wit01c, Wit01d]. Talk [Ano83f, Dav77b, Fel82h, Fri69, Hun03, Isa92k, Joh05, Kal82h, Lei74, Opp63e, Pan07a, Ano92s, Bar80c, Fel79h, Fly98c, LKWR89, Zag91b, Tic94]. talked [Ano90-61]. Talking [Gru04, Mar66b, Sol91, Bul08r, Ano98n, Yor88]. Talks [Ano49x, Ore60a, Ark89b, Bor86b, Dea89, Gar84c, Han98b, Isa85e, Isa89m, Mia16, Rob75b, Was97, Zam95c]. Tall [Gus01, Rot95h]. Tamerlane [Ano94-28]. Taming [Ano90-35, Ros61]. Tamplin [Cow71, Pau70, TG70, TB70, Tot70]. Tanaka [Joh97a]. Tang [Tra72b]. Tangled [All88, Rot94f]. Tangles [Dre99]. Tank [Ano93-37]. Tankers [Jas81]. Tanks [Ano93-38, FLLW91b, Rot95a, Zag94c, Gor73]. Tanner [Saw81b]. Tantalizing [Ano78m]. Tanzania [Nye79]. Taoka [Ano90p]. Tap [Aud65, Sim60b, Kil57]. tape [Ano91-59]. Tapes [Leb91b]. Tapinos [Rea80]. taps [Ano91-31]. Tarantula [And71]. Target [Mas08, Opp02b, Pos00, Al 15, Bin15, Kul88, Mak95a, Mil91, Moo96l, Ulg15, Row88]. Targeted [Nix10, SL13, Len20]. Targeting [Ark00g, Ark01e, Goo03, Isa90v, Mak03, Pik06, Kin78b]. Targets [Kri97, Pow74, Dav79c, Gor95c, Isa96d, Rot93n, Tob12, SC97]. tarnishes [Eps87a]. Tarter [Bul07b]. Task [Bel46, Fra47, Gou48, Gou55, Rab51k, Gol88e, GK21, NH08b, Wei87]. Tasks [Dea54b, Tsi89c]. TASMania [NA89o, NA92b]. Tass [Sal63b]. taste [Gra81a]. Tastes [Bar63]. tasty [Ano95n]. tat [Ano93v, Ano94n, Ano95m, Bar10b]. tatski [Ano96n]. Tatyana [Kap91a]. Tauber [For80]. Taubman [Leb91b]. Tax [Pay70, Rab59l, Ram70, Sta62a, Tau85, Cle11, Woo98a]. taxation [Isa96b]. Taxes [Eis68, Gol72b, vHWD76]. taxing [Ano92q]. taxonomy [Bye87]. taxpayers [Bal86b]. Taylor [Dow82d, Hal93a, Tur54, DeN74a, Har60a, Fol74]. Tea [Oli66]. Teach [Cam95b, Fla65, Joh66, She65, Wal82]. Teach-In [She65]. Teach-Ins [Fla65, Joh66, Wal82]. teachable [Rap11]. Teacher [Ano88e, Orm57]. Teacher-Designed [Ano88e]. Teachers [Joh83a]. teaches [Mec21a]. Teaching [Ano52g, Ano96v, Ano96w, Her79, Men64, Mul67, Rog88b, Sch58b, Som84d, La 88b, Rin84c, Whe84b]. Team [Ano56h, Ano99c, Wit01a, Mor77, Wit01c, Ano93-31, CP93a, TB23]. Team-B [CP93a]. teams [Kib12]. Tear [Gol70a, Mes71]. Tearing [Rot93r]. tears [Chu94a]. tech [Dro22a, Dro22c, Eav17b, Hor14, Isa91h, Kel98a, MW91, RS88a, Wal90d]. Technical [A.51, Abb79, ABKS10, And54, Ano45p, Ano53-45, Ano54-31, Ano60c, Dan54b, Fre84a, Hei47, Hur78a, Ing60a, Khr57, Kin54, Kri85, Leo65, Lev70, Lev72a, Mal54, Nad66, Rev68, Sei71a, Tel73, VT73, Wad58, Woh54, dF80, Bal82, Kel10, KR84, Wio79b, Ano60a, Ano60d, Ben71b, FK71]. Technics [Aue78]. Techniques [Ano60a, Eis90, Hic22, Ric98a]. Techno [Mar98c, Mil69b, Moo18]. Techno-babble [Mar98c]. Techno-Corporate [Mil69b]. Techno-fixing [Moo18]. Technocracy [She74b]. Technocratic [dF79d, dJ64]. Technocrats [SF77]. Technological [AP80, Ano69p, Ano86-28, BAB^{+}46, Gan83b, Gol69a, Gre67c, Gus85, Sal64a, Sch79b, Stu73, War68, Wel60, Zah69, dF79d, Bet85, Bro79d, Bro79e, Bug78, Der85, Dha16, Huq16, Jon84, Led85, Mec19g, Ste85b, SW16, Ven79, Rab72f, Wil81a]. Technologies [Bar61a, MF10, Tsi72a, vHW75, All14a, All15a, Dro22e, Hec22, JFS20, MA17, MvE13, Che88, Per87b, Tuc77]. Technology [Act09, Ano69r, Ano86l, Aue78, Aur71g, Bar60, Baz83, Ben71b, Bla57, Bla61a, Bro58, Bro71, Cal64, Dow92, Fel66c, Fel73a, Fel76h, FS84, Fin84, Gof71, Gou75b, Her88b, Hoa69, Hoe09, Hol74, Hym71, Joh60a, Kan69, Kre85b, Kri75, Lap57, Lar67, LEH^{+}71, Lon73, M.84, Mac68b, Mar98c, Mec63, Men68, Mil85d, Moc56, Mor79, Nel57, Net69, Orl84, Pas74, PO16, Pla73, PF75, Rab59f, Rab64f, Rod69, Sab72, Sei81, She56, Shu72b, Shu72a, Sim84c, Sin81, Smi71, Spo69, Sta58, Sut84, Tae51, Tel73, Tod78b, Tuv53, Voh71c, Wei66b, Wie62, Woh64, dF79g, dH75, AT92, Ahm80, All16, Bau84, BHH11, Bri80a, Cas78b, Coh86, Eve85, FP17, Haf85, Hal84, Her88c, HV18, Kal77a, Kow77, Lay79, Lif13, Lin86, Man87, Mec23c, Mec24, MK79, Mil81, MS77c]. technology [Mor83a, Myn11, Nye77, Pan81, Pat80e, Pol76, Pra22, Sch84b, Shu76, Smi77a, Spe89c, Ste76a, Stu87, Tir85, Tir86, Uni78, WKK23, Wei80b, Wio79a, Wri77, YG77a, Zin53, Ano78n, Ano81f, Ano85d, Dow80e, Pla57b, PF75, Tho80g, Tri71, Gil83, Gol78a, Hir88, Hor77, Mil88, Nor88, O’C87, PF75, Saw81a, Sie85, Sil79a, Sne57, Diw79, Pow91b, Ros68a, Sag86, Sol80a, Ste89e, Ste76a]. technopolis [Lat79]. Technostrategic [BB87]. Tectonic [Ano97k]. Ted [Mec23b]. Teeth [Rab98, Kib12]. Teething [Voh71a]. Teil [For80]. Tektite [Rig70, Kob70, RC70]. Telecommunication [Smi71]. Telepathy [Ras55]. Telephone [Ano78p]. Telescope [Coh69, Sch57a, Fel81k]. Telescopes [Spi61]. Television [DL78, Dow81d]. Tell [EW06, Mil03, Say84, Wei94a, Ano90-43, Ano98l, Rot92r, San91, Wal89c, Wu22, Cas77f]. Teller [Ano05b, Her90b, Mor62b, Pal02, Zac76, vH83b, Ano87-40, Ano89r, Ano90-31, Ano95i, Ano02g, BO76, Bul83, Cun88, Flo88, Goo04a, Goo04b, Har10, Her90b, Hir88, Mag88a, McM03, NA91m, Ore58a, Pal02, Pei80, Yor76b, Yor89, vH83a, Ber78a, Sch80c, Han92b]. Telling [Ahe90, Ano96o, MT96]. tells [Bon90]. telltale [Flo88]. Telstar [Lew62b]. Temelin [Wes99, Wes00a]. Temperaments [Gen84a]. temperature [Ram16]. Temple [Pow88]. Templer [Dro20b]. temporarily [Ust93e]. Tempting [Buk02]. Ten [Ack90a, All53, Ano63m, Ano93x, Ano96-28, Deu67, Gui65, Kel84, Lew56, Rab52c, Rab56i, Rot94a, Wur93d, Zag95e, Ano93-27, Wit95a, Nei82a, Rab20]. Ten-hut [Ano93x]. tenable [Dud76]. Tennessee [NA91e]. Tenney [Rus51a]. Tension [Fle58, Top58b, Cah94]. Tensions [Chi59c, Ima93a, Pif18, Puz93b, DM24c, Kor17, MFN17, dTdLF24]. tentative [Ano85j]. Tenth [Rab56k, Rab56l, Rab56j, Tur52, Men96b, A.62]. tenuous [Spi83, LLS83, Roc83b]. Tenure [Ash57, Gla56a, Din69]. Ter [Dro22i]. Ter-Mikaelian [Dro22i]. Teraflopping [Moo97d]. teraflops [Ano97i]. Term [Ant74, E^{+}84, Kri85, Lov74c, Mei51c, DM18, Due21, Ial18, Mes97, Nel12, WvH86, vHC86, Mei51a]. Terminals [Sha90]. Terminates [Cle01]. Termination [Bru57, Rot79a, Van88, Tho87]. terminology [Ano86-34]. Terms [Ano48q, Ano48c, Ano55n, Bol88c, Ano86-33, Ano87-34, Bro84b, All55b]. Terra [Lew69b]. Terrain [Ber93b]. Terri [Dow83b]. terrible [Isa92b]. Terris [Pop64, Ber84b]. Territories [Soh61, Lev70]. Terror [Cho02, Dis04, Isa02f, Isa03e, Mas13, Mur60, Nad98c, Opp02b, Rey04, Ric02, Rot03e, Ano94o, Ano97k, Yor76a, Mac84b, Nei82a, Sei62]. Terrorism [Ano87o, CPV01, Col00, Har05, Hav03, Isa01e, Llu99, ML09, ML10, Sch05a, Wit03b, BB20, Coh76c, GCV16, Hec16, Hol76b, Int86a, Khr10, KSG15, Wei15, Nei82a, Veg98b, Fly98a, Eld92]. terrorist [AB18, Bla11, Kri87]. Terrorists [Gru04, Gyo77, Kri75, Wil75b, HMR86]. Terry [Pee78]. Test [AM21, Ano46h, Ano46k, Ano46g, Ano46-64, Ano52i, Ano53-38, Ano56i, Ano56m, Ano58e, Ano60r, Ano61p, Ano63n, Ano63u, Ano63r, Ano63s, Ano63t, Ano79e, Ano90n, Ark86h, AIL^{+}62, Bec64, Bor53, Bro85f, Cha88b, CP99, Dav60, Eps88b, Eps91, Fed58, Foz61k, Foz62i, Gar74, Gea89, Ham58, Hla65, Hum59a, Ing59a, Ing60a, Ing60b, Ing63b, Ing64, Ing79b, Isa98e, Isa98f, Isa99f, Joh95, Kan69, Kas21, KK63, Kia93, Kim01, Kol64, Lal67b, Lay07, Lib59a, Lon82, Mar83e, Mat83a, McN63a, Moo97p, Ore59, Ore60b, Pif24, Pre63, PE89, Rab56a, Rab57d, Rab59m, Rab63j, Ros59, San63, Sch01c, Sim59d, Spi59, Szi63c, Tam74, Tur60, Ure61, Wes58, Wie82b, Wil83, You61, Zam92d, Zam98, Zei53, Zer13, vH91, Alb87b, Ano53o, Ano86-30, Ano95n, Blu23, Bro88b]. test [CSF95, Cle89, DM81b, DM85, DM86, Div86, Dre77, DT91, DvH91, Eps86, Eve85, Eve88a, Eve88b, Fet87, Gaf86, GH83, Han90d, Isa90s, Jeh60b, Joh96b, Kat16, Kid91, Lee24, McK86, ML88, NA87j, NA90j, NA90k, NA91g, NA98a, Pod24, Pom77, Pom79, Rot96i, SL87, Sip86, Tov89, Wit82, Wol88, Yor81, ZK94, Ano93-30, ML09, Nel09]. Test-Ban [Bec64, Lal67b, Ore60b, ZK94]. Testifies [Ano64l, McN67b]. Testify [Etz64]. Testimony [Ano49t, Ano56l, HPL^{+}49, Ric81, vN57]. Testing [Ack86, All57k, Ano60m, Ano60o, Ano86-29, Ano98m, Ben69, Ber86c, Bro74, Dan90b, Fel62a, Fel62b, FM01, Foz62i, Gil69b, Gil69a, Ham91c, Isa90w, Isa92n, JCF02, Kal62a, Kal60, Kal61, Los57, Led71a, Lor01c, McC57, O’B69b, O’B69a, Ore58b, Rot04g, Rub62, SD56, Wig58, vHZC93, Ano85q, Dan91, DM84, DM24c, Eps87b, ES89, Eps90a, Eps90c, Gol88f, Hop77, Kal85c, Lan89e, Mey86, MN96, Mil86a, Nel92c, NA90p, NA98e, Owe87, PE24, Pin24, Rot93d, Ste60c, Sut95, ZC92, Zam92a, Zam92c, Zhe90b]. Tests [Adv60b, Ato54b, And60, Ano46i, Ano46h, Ano46k, Ano46g, Ano46j, Ano46q, Ano46-99, Ano50x, Ano50y, Ano51m, Ano55r, Ano57k, Ano59i, Ano59j, Ano61v, Ano62-27, Ano88e, Arn54b, Bae60, Bal60a, BBG^{+}85, DuB46c, Fed59a, Fed59b, Gre58, Hum60, Ing54a, Ing57b, Ken60a, Kis54, Lau46, Lib58, Mac54a, McE88, Neu58b, Nor55, NA88e, NA90e, NA93h, NA96c, O’B48, Oga58, Pau62, Rab58d, Rab59n, Ros59, Sch57b, Ste69e, Str54b, Thi58, Tur53, Zei52, Alb98c, Bir89, Dun85, Fre93a, Gub15, IS92, Ide93, Isa92e, Jeh60b, Kin78a, Mit58, Nag15, NCA88a, NCA88b, NA89d, NA89h, NA89g, NA91f, NA92g, NA94b, NA95c, NA98b, NA98c, Ruf90, Ste86a, Wes79a, Wri97, Zha15c, dTdLF24, vHJ96, Ano60d, Pet81, Lap59b]. Tet [Sox80]. Texas [Wyl56, Kur79, Gre87]. Text [Ano50-47]. Textbook [dF80]. th [Ano60f, Gol66, Ikl60, Lak64, PK63]. Thackray [dF79f]. Thalidomide [Gol73b]. Thanks [Ano75e, Ano97o, Ano98p, Pug05, Sco95, Jef94, Mar90, NA89s, Uhl95a]. Thar [Sha69]. Thatcher [Pat90a]. Thaw [Win90, Cra86b, Cum92b]. Thaxton [Dow81e]. Thayer [Sch68f]. Theater [Ark01f, Lin84a, BH80, FR98, Gho11, Pai83c, PP83]. Theft [DeN74a, Fra75]. Their [Ass53, Aeb48, And78a, Ano57h, Ano57k, Ber93b, Dil57, EBU^{+}52, Hal75a, JCKG^{+}50, Lap56a, Mul47, Nea02, Osb54, Pip55, Rus56b, Sha90, Ano93u, BW18, CKN^{+}94, Fit11, Har87c, HSJH18, Pol91b, Tel47b, Wei85b]. theirs [Ark89a]. Them [Ano58l, HvH75, Siv75, Ano95n, Ano97k, Ark95d, Dou24, Fie19, Isa91r, NA91m, All59l]. Theme [Aue78, Soh62, Ano94k, Tur77]. themes [Tod78a]. Themselves [Ano53-40, Lan90h]. Theodor [Has71]. Theodore [Ano86k, DeN74a, Fol74, Gre68b, Hym68, McL49, Sal63b, Ste91a, Sto56b]. Theodosius [Cro67]. Theology [McL66, Pop52, Ima93b]. Theoretical [Wei55c, Wol55]. Theories [Lit54, O’K69, DS15, Sin92]. Theory [Ano56b, Ano57j, Arm68, Buc50, Cah94, Cra83b, For80, Fre09, Jas81, Kol07, Mic74a, Nei82a, Koe77, Kra94, Ros87b, Cha75a, Was62]. Therapy [Sac72a, NA91e]. There [Lan01, Mur55b, Pie19, Ros75, Tay47, Ano98k, Coa88, Coh98, Gol84b, Hic22, Hus24, Isa94h, NA90p, NA91m, Pat80b, Sch16, Set16, Sim16, Wil15, FJ12, Say84]. therein [Due21]. Thermal [Gus70]. Thermodynamics [Ber72c, GB72]. Thermonuclear [Ano53-38, Art71, Cla61, Kur56, Lep62, Mar61d, Tel56]. These [Zag94c, Gol96b, Nor94b, Rot83]. Thesis [Dys69d, Jen70, Lew70f, Rab70b]. They’re [Jag95, Ano97k, Blo95]. they’ve [Ano90-54, Fra90a]. Thick [Bro55a]. Thiermann [Say84]. Thin [O’L75, All14b, Kim95, Woo91]. Thing [Cum01a, Uhl01b, Wil73c, Ano93z, Dro24f, Fel80d, Moo92g, Sho92b]. Things [Ark01d, Kat61b, Rab59e, Rot01h, Ano90-34, Dro21a, HV18, McK25a]. Think [Ano90-31, Bal93, Ela93, Els90, Kat61b, Ano92v, Dah77, Dro22g, HSV19, Mar77, Mil78, Gor73]. Thinkers [Ran85, Kap82]. Thinking [Ano47-28, Boy10, Bro47a, Bro47b, Bur76, Cav07, CS17, Ing55b, Low93, MZM01, McK07, MK51, Nas83, Rab47e, Sch71b, Ack87h, Ano95k, Bar90a, Deu91, Gri87a, GT14, Isa90n, Pat79f, Pra89c, Sea89b, Sel77, She85a, SWC^{+}84, Ano85c, Dow82g, SWC^{+}84, Cra88, Dia83, Fri62]. thinly [Smi97]. Third [Ano48-39, Ano58l, Ano58k, Ano86i, Bea89, Bus85, Eas85, Fal61a, Fis83, For75, Gol69b, Gou78b, Hun87, Joh89, Kal70, La 66, Ler80a, MvHKN01, Med76, Mer90, Pat82a, Pow90a, Rab58f, Rao77, Tel76, Tod79, Val86, Dw77, Dje74, KE17, Ack88l, Ano90-64, Bro58, Con90a, Cra84f, Dha90, Fed59b, Fed59c, FKM87, Got77, Gou75b, Har87c, Hay78, Hei68, Kal83a, Kar88c, Kar88d, Kla84, Kla90, Lev70, LGW92, Mah59b, Mah59a, Mak76, Pow89a, Rav84, Red78a, Red78b, Rev92, SR80, Sal88, Say84, Top58b, Ven79, Wei58, Wei69a, Win90, Hun87]. Third-World [Kar88c]. Thirring [Bor52a, Wei48]. Thirsty [Wri84]. Thirteenth [Tur53]. Thirty [Ano89y, Bec75, Dow82i, Nel75, Sti83, Tsi82d, Yor75b, Yor82]. Thirty-seven [Dow82i, Tsi82d, Yor82]. Thomas [Abr74, Ano78d, Ano78n, Ano86i, Ano89g, Ber73c, BGR61, Bro54, Bul67, Cal84a, Dav69, Dia83, Gol88a, Han98a, Hun87, Met83, Sei71b, Sei73, She61, Sta64, Swe93, Tho80c, Ano12f, BCH^{+}48, Kat57a, Sci51, Sei73, Bul08s, Sto59b, Wri47a]. Thompson [Ano86j, Gus83, Wig62a]. Thoms [Ste86c]. Thorium [Lov79, Nel12, Bal93, DLF79, Fri97, vHWL79]. Thornton [Bro61, Hun87]. Thoroughly [Pap88]. THORP [Mos93a]. Thorwald [Cre68]. Those [Ark93a, The89c, Thi58, Ano89k, Ano90v, MRH16]. Thou [Eps90a]. Though [Roc81]. Thought [For86, Mor48b, Rob91b, dF79h, Aue78]. Thoughts [Bre67, Dys62, Dys63c, Gom61a, Gom61b, McN59, Mor62a, Rab58c, Rab62g, Rab63c, Rab64d, Rab65f, Rab66d, Rab67g, Rab68b, Rab69e, Rab70d, Rab71d, Rab72d, Rab72c, Rab72g, Shi54a, Ste53, Wei59a, Ano90-31, Day77i, EG24, Fel76f, Gom80, Med76]. Thousand [Ano46a, Bet58, LH70, Ano97j, Roe91].

Threat [Ant69, Arn54a, BB10, CPV01, DeN76b, Die91, Dow82f, Fin54b, Fot60, Gar08, Gov93, Gre67a, Hec06, Jac56, Leg56, Lug05, Sch05g, Sha90, Bul01c, Bul07d, Str06g, TS99, Zha99b, AH20, Ano97n, Arn97b, BDE^{+}77, Ban11, Ber81, Bla11, Bre13, Cho81, Eps90c, Eva87a, Kre86a, Kre87b, Lin19, MDB81, Mec17c, Mec17b, Mec17d, Mec18c, Mec19g, NA87i, PWS^{+}96, Pop19, RK19, Rot81b, Shu81, Smi16, Ste81a, Wan96, Ano87o]. threaten [LP85a]. threatens [Bur90b, Eri88a, Eri88b, Tir86]. Threats [Ing68b, NK06d, RH14, Tam71b, CK98, Cro17d, Gro15, Kir55, Kos19, Mec23b, Mey85, Rus92a, Smi21, Sto18a, Vog13, VA11]. Three [Ano80k, Bak75, Car76, Cha79a, Cof65b, DeB81a, Fel74b, Fel84e, Fri48b, Fry65, Gen84a, Ing59b, Isa81a, Isa93k, Jen70, Lar65, Lew70f, Obe97b, Oli89, Rab50d, Rab70b, Sti83, Voh70, WKW57, Wyl57, YTS61, Zim81a, Alb98c, Ano96o, Ano98n, Ano12b, Ark97g, Ash81, Day77i, Fel82a, Gam94, Gus11, Lin79, NA89i, PE89, vH91, Atl47, Col79a, Fai79, Fel79l, Fri11, Gil80, HK13, KKHK79, Man04, Sti83, Wal83a, Wal85b, Mas84]. three-say [Gam94]. Threshold [CLR75, CB06, Oft71, Rab70d, Rot96i, Mar83e, Wil83]. Thrift [Ber72c]. Thrive [Mot51]. thrives [Rab85]. thriving [Kla88]. Throw [She74a, Str90]. Throw-Away [She74a]. Throwing [Nie66a]. Throws [Woo99b]. Thucydides [Leb79, Mec22b]. Thunberg [Dro23d]. Thunder [Kri85]. thunk [Ano98p, Isa88a]. Thwaites [Dro20a]. thwart [BB20, Isa86f]. Thwarted [Cic74]. thwarts [Ark87a]. thy [Ano95n]. thyroid [vH80]. Thyself [Ben69]. Tibet [Ahm00c]. tic [Koe22]. Ticket [Fel73c]. Ticking [Hal93b]. Ticks [Kra10]. tidal [Pai82a]. Tide [CSG71, LSP95a, Mic74a, Rab71c, Smi72b, Isa91n]. Tied [Mei06]. Ties [Pow90a, Uri94c]. tiff [Lan89h]. tiger [Boo22, CK20]. Tigers [Ano91-57, Ano90v, Ano90-28]. Tight [Ano68k, Moo95j]. Tighten [Wei91b]. tightened [Nov77b]. Tightening [Spu73, NA90l, Rot77]. tightrope [Fel82j]. Tikho} [Ano93-27]. till [AH92f]. Tim [Bri86, Dow80i]. Time [Abb90, And74b, And74a, Ano86h, Ano95x, AB05, Bai75b, Bai75a, Ben01b, Bla97, But97a, Cas77c, CFM^{+}10, Con58, Dre82b, Eps77d, Fel75b, Fel80i, Fel92, Fit75, Fri69, Gre67d, Gub15, HI05b, Hen68, HNR12, Hol94f, Hud76, Hue04, Isa89d, Isa00d, Jac74e, Joh91, Kam74, Lor77, Man74, Mar87b, McD74a, MM71, Mie74, Nad80, Nag15, Nye77, O’B69b, Opp47a, Pai05, Qua80, Rab62e, Rab62j, RD91, Ros62, Rot04g, Sea61, Sie05h, SD15, Spi75, Str07f, Tod79, Tsi78a, Tsi90, Wat74, Wei94c, Wil74a, Wil75f, Wil75c, Wil75d, Win53, Woo00d, Woo00e, Zha15c, dH75, Ack87b, Ano93s, Ano96o, Ano98l, Ano98o, Bac80, Bai20, Bra17, BHK^{+}21, Car97, Dro21c, Fel77a, Gor92, Hal93b, HI05a, Hol77, Hus24, Isa90z, Lac90, Lau91b, Man95, McA84, MN96, Moo91d, MT96, Mye95]. time [Pai83e, PC92b, Rev21, Rob20, Roc97a, Rot98c, Sch16, Sch07c, Set16, Sim16, SCHD20, Vos88, Wur93d, Zam96b, Ano98h, She61, Eli85, Mul72, dF80]. Timely [Fis89, Moo91f]. Timerbaev [PT87]. Times [Ber78a, Cle71, Dow81e, Har79b, McM52, Mea93, Olt52, Opp56, Ros52, Swe88, Web03, Zag94a, BO76, Mec18a, Mec21e, Opp20, Rot97i, Bre93, Ada67, Ano78l, Bro54, Bre93]. Timid [McG67]. Timing [Str06f, Ano90-31, Ano91w, Zam96a]. Timmerman [Har92b]. Timor [Sut97]. Timothy [Cra84e, Van03, You66b]. Tinbergen [Lyt79]. tinfoil [Str90]. tinted [Dor85b]. Tinynukes [NA92o]. tipping [Dro25b, Dro25a, KGS25, McK25b, NLS25, Spe18]. tired [Sch91]. Tirman [Ano86i, Lac85, Sie85, Uhl98a]. Tismaneanu [Gol93d]. Tit [Ano93v, Ano94n, Ano95m, Bar10b]. Tit-for-tat [Bar10b]. Titan [Bel55, Fel80j]. Titanic} [Joh95]. Titanizing [MS61]. Titans [FP90c, NA87g]. Titelbaum [Ekl80]. Tito [PMS00]. Titski [Ano96n]. Titterton [Fer57a]. Tiz [LSP95b]. Tiz-moes [LSP95b]. Tizard [Lak66]. TKO [MH89a]. Tlatelolco [Gol83b]. TMI [Sox80, Ano80k]. tobacco [Whe90]. Today [Ano56f, Ano66b, Bro57a, Col03, Dob52, Dow80j, Fed72, Hor71, Kah50, Kli52, Lon52c, Lov74b, NK07f, The89d, War63, Wei56a, Ano98r, Cor84, Cra83c, Dro23a, MD18, MSW20, Rot81b, Tud21, Wu22, Nor64, Rab72a, Sch15a, Ano91r, Gol88a, Kus83]. Together [Ano46-104, CRB^{+}09, Rab62k, McK23a]. token [Rot95b]. Tokyo [DM24b, Jac58]. Told [Gla75a, Rab62h, Ano94m, Sha91b]. Toledano [Gre64a]. Toledo [Sal80c]. Tolerance [But00, Ken07, LaF90b, Rot96i]. toll [Bar77b]. Tom [Ano87q, Dow80j, Dow82b, Dow83e, Jon87, Rot97b, Sto59b, Ano14g, Cro18, Dro21d, Fie19, Lor01d, Mec21d]. Tomahawk [Ark84e, Arn90, Arn91, KA91, NA87d, NA87j, NA90p]. Tomahawks [Arn92a, NA89f]. Tomorrow [Bro57a, CCH72, Col03, Cro65c, Dow80j, Fra72, Hib96b, Mag73, Maz69, McP65, NK07f, Ano96o, Ano98r, Gei17, MD18, Ano78n, Smi59b, Mes64a]. ton [AZL90a, Han90d, Moo97n]. tongue [NB78]. tongues [Isa90l]. Tony [Rom90]. Too [AH02, Ano90-31, Cho03, Cut75, Eas00, Ger75, Gro06b, Hub57, Hun00, Ing77, Kre93, Lau60, Low76, Mar02b, Moo93j, Moo95p, Myr68, Pit52a, Pit52b, Sau74, Sau01, Sim59e, Smi94b, Wri04, Ano92t, Ano92q, Ano93s, Ano95n, Ano96o, Ark86c, Ark89a, FR97b, Gar97c, Gei16a, Gey94, GH94, Gub15, Jon91b, Kal95b, Lev89, Lun79, Mar02a, McK25b, Nag15, Nor81, Pai83g, Pan22, Rot98d, Wit95a, Zan91, Zha15c]. Too-Modest [Gro06b]. took [Ano89r]. Tool [Fre72a, She77, GK85, SD17]. Tools [Fla90, Wan05, Zah06, Alm85, Eav17b, MA17]. tooth [AM21]. toothless [CK20]. Top [Aud65, GL95, Gor60, Han92b, Nor04, Sim60b, Spo62, Sti62, Ano87z, Kil57, Kla77, Riv96]. Top-Secret [Han92b]. Topaz [Rot96c, Tar93]. Topchiev [Rab63a]. topic [Ano14b]. torium [Zam92d]. Tornado [Wyl56]. Tornados [Ano53-46]. Tortoise [Cal60a]. Tortured [Mic74a]. Toshi [Mes83a]. Toshiba [Ano87-41]. totally [Ano98m]. Tottering [Joh99b, Smi92d]. touch [Isa82d, NA92k]. Tough [Jac74d, Lan10, Wes99, Gad89, Gra98]. Tour [Ric71]. tourist [Pre92]. Tourists [Ano97i, Hyd91]. tours [Sto90]. Tow [Sim85]. Tower [Ant76, Pei69, Tuc72, Wil76b]. Towers [Jac74a]. Towing [Lon74, WC73]. Town [Lan90l, Mol69]. Townley [Lal70]. Townsend [All55a]. Toxic [Abb84a, Gov93, Tho80c, Tho80b]. Toxicity [Day77d, Eds76, NE76]. Toxin [McL11]. toy [Koe77, Moo91h, Ano97k]. Toying [Zem06]. Toynbee [Rab60b]. Toys [Ano98p, Bar78c]. Trabi [Ano96k]. traced [Fre77b]. Tracers [Hev47, Wol55]. Tracing [Arn53]. Track [BST02, Cal97, DeV92]. Tracking [Ano11f, GP92a, Ste90a]. tracks [Gar97b, Kem97]. tracts [Swo86]. Tracy [Dro17]. Trade [Kas64, Sol95, Cle11, Har90d, Hof88, Kla79, McC88b, Nye79, SSK10, Veg97, WL84, She79, Uhl98a, Sol80a]. tradeoffs [Ack90h]. trades [Kla88]. trading [Wig88]. Tradition [Hei73, Ack84]. traditional [Fri11, KM18]. Traffick [Wie06]. trafficking [Epp20]. tragedies [Wur93d]. Tragedy [Sch02b, Bal84a, BS05, Lif23]. tragic [AK81, Ano94l, Hen78a]. trail [O’N95, vMC98]. Train [Sie04a]. Trainer [Gil83]. Training [Ano46-28, Ano49-33, Kat52c]. trains [NA89m]. trans [Har76b]. trans-science [Har76b]. transatlantic [Pos21]. Transcaucasus [Sar89, Uri94c]. Transcend [Lov74b]. transcends [LL16]. Transcript [Ano54z]. Transfer [Goo86, Gou75b, Sag86]. transfermium [Rot95i]. Transferring [Rot03e]. transfers [Ano87v, Kla84]. transform [Hil12]. Transformation [Mei70, PD12]. Transformed [Moo98b]. Transforming [Bel20a]. Transient [Mid73]. Transit [Lew69b]. Transition [Ano78m, Ber75c, Isa88a, Lew70b, Owe79, Par66b, Spi66c, Sta85, Uhl95a, You66b, Fri21, Gor81, Not15, Sch13b, Sob21, Wal21b, Dre82b]. Transitional [Sel69b]. Translated [Ber79, Bar67b, Cog82, Ekl80, For80, Leb91b, Rab69b, Rab72a, Rab73d, Ran85, Sha54b, Sig85b, Ten75, Tho80g, Tod81, Tra71, Wri63a]. Translation [Kul06, Rit67]. Transmission [Mor73, YY74]. transmitted [BAB^{+}46]. Transmutation [Gib91b, MZM01, PG91]. Transmutation-related [Gib91b]. Transnuklear [Cha89d]. Transparency [Buk03, Bul08b, DS12]. transparent [Cot19, GG11a]. Transport [Hoh66, Lor02]. Transportation [CCH72, Hir73, Wal84, vB69]. Transporting [Ano78n]. Trap [Gal63, Boy05, Mec22b, Wei82a]. Trapped [Ses83, Ano97i]. Trash [ADK91, Web02, Azh93, You92]. Trashing [Nad98d]. trauma [McK23b]. Travel [Apt64, Bor66, Lal66i, Tol58, Ano95j, Khr97, Nea78a, Nea78b]. Traveled [Lan96]. Traveler [Van03, MK02a]. Travels [Old69]. Treading [Isa93l]. Treasure [SM14]. Treat [Spa95f, Bar80a, You85a, vV23, Coh53b]. Treaties [Ano62q, Ano66n, Joh99b, Kre99, McG84, Ano89c, Bil85, Dor90, Kre97, Loe89, Mat83a, Mey18, Ros94, Sco89, Wit82, Ano94s]. treating [Isa90w]. Treatment [AME50, Guz52, Mos52, Rui68, Val52, dO52, Sha90, Ste77c]. Treaty [Bec64, CLR75, Cle01, CP99, Gor67a, Hol94f, Ing64, Isa97e, Joh68, Mei01, Que69, Riv96, SS90a, Spi59, SK03, Tho90, War91, Wei61, Zer13, Alb93e, Ano86r, CAM20, DeV92, Duf87, Far87, Fel80l, FLLW91a, Gol84a, Isa87e, Isa87d, Isa95e, Isa96d, Joh95, Kap90c, KT97, Kyl97, Loe87, Mec20a, Mei98, Moo97p, Mül90a, Pai83a, Por19, Rot92l, San90, Sco90, Sim11, Smi92d, Smi98, Tho87, Wri91, Wri97, Ack87f, AC85, AM21, JFB19, Dan70, Eps81, Fel67a, Fel68a, Fel68c, Hal75b, IK21, Ima69, Kas21, Kre86a, Lev85, Lev87, LP85a, Mar83e, Men96b, Nel09, Pai02, Pif24, Que67, RKK22, Sut96b, War88, Wil83, Fre88]. treaty-making [Isa95e]. Tredici [Day88]. Tree [Edm74]. Trees [Ros61, Coh87, Fie19, Isa94h]. Trek [Smi87b, Tho80d, Gom87]. Trek/Star [Smi87b]. tremors [Ler80b]. Trenches [Coo71, Ano96l]. Trend [Net71, Shi47d]. Trends [Ano70f, Ano70g, Ano71h, Ano87-42, BHH11, Kob70, Maz69, Rot62, Tod79, Com75g, Moo98a, Pow82]. trendy [Art85]. trespass [Hun90]. Trever [You84b, You84b]. Trevor [Don90]. Triad [Khr00, Ric09]. Trial [Ano50r, Bin67, Che50, Gro56b, Sim54b, BC96, Mar78b, SG69a, SG71]. Trials [Ano51b, Dow81i, Rab51a, Ano78i, Cha78a]. triangle [Fel79k, Yud83]. triangles [Cas77a]. Tribe [Ber77b]. Tribute [Fel84g]. Tricks [Zel76b]. tricky [Rot95e]. Trident [Ano75d, Ano80r, Ark84f, Mil02, NA87d, NA89m, NA90c, NA90q, NA91e, RD80, Ros75, Ros89b, Sha00]. Tridents [NA89j]. Tried [Ing55a, PG91]. Tries [AK01, Bro88a, Eav18, Isa87c, Isa88f, Isa90m]. Trieste [Gom75c, Sal80c]. trigger [Arn92c, Mar89f]. trillion [CP93a, Nuc95, Sch96]. Trim [IK01, Cai90]. Trinity [Wea82, And74a, Bai75b, Blu23, Kis80, Fel66b, Mos92]. Trip [Jac73d, Van03]. triplicate [Ark86f]. Tristram [Saw79, Cof62a, Cof62b, Cof62c, Cof62d]. Tritium [Buk97c, Fer99, Qui84, AT88, AB89, Ano84j, HK96, Kal95b, Lan89h, LH88, Mar90, Mil89b, Rei91, Rei93, Rot97k, Sto97]. Triumph [MR07, BS05]. Triumphalism [Uhl03a]. triumphs [Dre91]. Trivial [Wag75]. Trojan [Com76b]. Troops [Day75f, Spa95f]. Tropical [Mar68, Mun54, NN80]. Trouble [Alb98d, Cab97, Fly00c, Hal75b, Har70, Joh00b, LC98b, LD00, MJ68, May08, NA87j, Sar89, Sch89c, Cha89b, Che97, Dou15, Gol23, Hor88b, The89a]. Troubled [Joh99b, Kre99, Mul65, Val97b]. Troubles [Voh71a, TM96a]. Troublesome [Moo95e, Ano88t]. Truce [Nei82a]. Trudeau [Smi85b]. True [BDF^{+}92, Cir05b, Har88c, Ano97k, PG91, Pra22]. Trugbild [Ano75g]. Truman [Bou02, Cap98, Mes84b, vL92, Ano45q, Ano46-77, Ano46-99, Ano48-34, Ano49-34, Ano50-39, Ano50-38, Ano52s, Ano53-41, Ano56g, Ano92r, Bak86, Ber84c, Bou02, Dow82h, Gol98, Mes85b, Moo03a, Mos85, Off02, Rab47l, SB84, Tru46, Tru47, Tru48a, Tru48b, Tru49, Tru50, Tru53]. Trumbull [Bru59]. Trump [Ald17, Cro17d, Kir17, Mec17d, Squ17b, WC17]. trumpet [Cas76, Har60a]. Trust [Ano81c, FM01, Hei89, Hic22, Joh79, Law88, Mei01, Moo93k, Pra06a, Wad58, Ano95m, Bec79, Bro86a, Dro23c]. Trustee [Hel77b]. Truth [Ano99z, Arn92a, Bri48b, Com75a, EW06, For75, Jac92, JCF02, Kin57, Mar91a, McN72, Mor83b, Sic96c, Sto55c, Uhl04b, Wal77, Dro24b, Fra47, Kat80a, Kel98b, Rot97c, Tho12]. Try [Ano90-48, Dre74, Lev96, Ano92y, Cla91, Man15]. trying [Dah77, Dro25c]. Trytten [Ano50-48]. Tse [LM76]. Tse-Tung [LM76]. Tsipis [Cal84a]. tsunamis [Mer24]. tu [Ano96k]. Tuan [Fra81]. tubes [McG90]. tucked [Alb94g]. Tucson [MS61]. Tug [Ahm03a, MW91]. Tugwell [Gou75a, You56]. tulips [Ano91-35]. tumbling [Isa98b]. tune [Ple91]. Tunes [Sch01b, Ano92m]. Tung [LM76]. Tunnel [You73]. Turbine [Cic74]. Turbines [SC75]. Turbulent [Rot97i]. Turco [Rob91b]. turf [NA87f]. turkey [Tay89, Ark85e, Kib03, Kib04, McK99, NA88m, NA90k]. Turki [Dro16b]. Turkish [Kib07, Ulg15]. Turkmenistan [Bus94]. Turn [AZL90b, Isa95f, Lor00b, MM98, Wad74, Dro23i, Kor17, NA90n, SD15]. Turnaround [Kib07]. Turned [Fri74b, HH75]. Turner [Mil85b, Eld92, Moo97b]. Turning [Abb04c, DeB81a, Dje75, Fry56d, Han75, Ing55b, Jac72a, Rab58e, Rab73g, Sti83, Cra86a]. turns [Joh94c, Pea85, Sok20]. Turok [Gri83a]. turtles [Bar90c]. tutorial [GJ84, Hol83c, Mar83b, Mor83c, Mor84c, Rom83, Sar83, Tsi83d, Tsi83e, Tsi83i]. TV [Hal93a, Azh93, Hal91, Lak78, Oze92, Rot92c]. TVA [And65, Gre67b, Fer99]. twelfth [Ano64p]. Twelve [Bri86, Qua80, Wil75d]. Twentieth [Bam85, Bro62b, Fra83, Gla76c, How65b, Mor62d, Pal02, Ste91a, Har10, Joh65c, Rip68, TS01]. Twentieth-Century [Gla76c, Pal02, TS01]. Twenty [Ano89z, Ano90-62, Bal79b, Bas80, Bri86, CK20, Mul65, Pat81c, Rab70e, SA62, Ber84b]. Twenty-first [CK20]. Twenty-Five [Rab70e, Ano89z, Ano90-62, Pat81c]. Twenty-Ninth [Bal79b]. Twenty-One [Bri86]. Twice [Dow82g, Mes85a]. Twilight [BS01, Gol96a]. Twin [Bar92a, Cav82, GC83]. Twinkle [Moo02d, Tar93]. twins [Flo88]. Twist [Fri59, She65]. Twisting [Mar02c]. Two [Amd65, Ano48q, Ano48c, Ano59e, Ano85s, Ano90-49, Ano92y, BB52, Ber49, Ber72b, Boo46, Bos51, Cha63b, Cum98, DS47b, DP53, DeV92, Dew53, Eps93, Foz63b, Gar97b, Ham91c, Har72a, HC87, Isa01b, Isa02a, Key66, Kre90, Lan61b, Lan90g, Loo50, Nas03b, NA89i, Opp02b, Ore59, Pop52, Pea92c, Pol49, Rab47m, Rab59m, Rab63b, Rab68f, Ric67, Rob75a, Ros59, Sal63b, SH95, Sha54b, Shi56b, Sim47, Sim65, SW80b, Ste83b, Sti61, Tel47a, Tel63, WN79, Wie48, Woo02, Yar80c, Alb98c, Ano86e, Ano91-35, Ano91-57, Ano97k, Ark95b, Bro88b, Cot22, Cra87, Eké21, Eve88b, Isa89a, Isa89c, Kem97, Kla19, Lym11, Smi21, Str62, Ano92q, Sha84a, Wei55b]. two-step [Isa89c]. Two-track [DeV92]. ty [Gus95]. Tydings [Ano50-45]. Type [All60h, Shu76]. Typical [MK03, Dro24d, Van03]. Typos [Rab56d]. tyrannize [Dou24]. Tyranny [Chr71, Gom68, Kas67, Sch24].

U [Hyn69, LaF86, Phi60, Rot93a]. U-2 [LaF86, Phi60]. U. [Iff67, Med47]. U.K. [Tur60]. U.N [Fin65, Hul98, MS77c, Mor79]. U.N. [Ano50q, Ano50-39, Ano56n, Ano66s, Bac58, Car49a, Cha72, Dal92a, Dea95, DFF^{+}63, Dor90, ER84, Eps88b, Eps92a, Err58, Fed58, Fed61, Fah91, Gol88e, Gol61c, Gol62c, Gro63, Hei89, Her88a, Jac62b, Jac63a, Jac64a, Ken63b, Kov68, Lal66h, Le 77, Lie46, Man66b, Mar65e, Moo95q, Pea88, Pom65, Ric75a, Rin66, Rob49, Rus92b, S.55a, Sch92b, Sch62c, Sco89, Sha78a, Sha78b, Sme76, Soc55, Sta62a, Sta62b, Wal50, Wat70, Wri47b, Wur91a, Wur91b, Wur99]. U [?]N.E.Mattern:1984:CPU. U.S [Ano84g, Ano86l, Ano88e, BBG^{+}85, Cal84a, Col84a, Cra83b, Dow88, Fra83, Gov93, Ham91c, Hun87, Kin80b, Kin86, Lei92, Mye87a, RS52, Roc81, Scl79, Som91, Ste86c, Tho84a, Tod78b, Woo82, dF79h, Kih47f]. U.S. [Adv50, Adv62, Ack88m, Ack88n, Ahm03a, Alb86, Ali57, ACC^{+}67, AS85, Ano46f, Ano46-78, Ano46-73, Ano47c, Ano47w, Ano47-27, Ano48-32, Ano48-41, Ano49-29, Ano49-39, Ano49-40, Ano50v, Ano50-35, Ano50-49, Ano51r, Ano52c, Ano52-27, Ano54q, Ano54u, Ano55s, Ano59j, Ano77n, Ano77m, Ano78o, Ano79n, Ano80m, Ano84a, Ano84p, Ano86v, Ano86-31, Ano87j, Ano87n, Ano87-30, Ano87-42, Ano87-44, Ano87-45, Ano87-43, Ano88s, Ano91-42, Ano91-59, Ano92-31, Ano99-29, Ano00o, Ark83d, ACH84, Bal85, Bar90a, Bar87, Bas80, Ber86a, Ber86b, Bet86, Bla85b, Bor87, Buc82, Bur87, Bur09, Cam90, Car84a, Cha85a, Cla74, CAH84, CAN88, Coh72, Coh92, CD80, Cra84a, Cra85, Dav57c, Dav79a, Ded66, Der85, DH86, Dow91, Dre51, DG07, Fai97, Fal89, Fei60, Fei61, FUvH85, Fly97, Gab01]. U.S. [Gol90c, Gol78b, Goo85b, Gro56a, Ham61c, HA85, Har90d, Har93, Har69b, Hua02, Ing52a, Ist67, JL74, Jon70, Kal87, Kat50a, Ken61b, Kes85a, Kih47f, Kin54, Kin57, Kom80, Lal64c, Lal64j, Lal66i, Lal67f, Lal67g, Lal67h, LD89, LeR50b, Lei90, Lev88, LP85a, Luc90b, MZ96, Man66b, Mar82a, Mar89f, Mat85, Mat84, McN67b, Mei01, Mer90, MS77a, Mor85c, Nea63, Nel09, Neu65, NH08b, NA87a, NA87h, NA87l, NCA88a, NA88s, NA88a, NA88r, NA88t, NA88q, NA88p, NA89q, NA89p, NA89r, NA90s, NA90r, NA91p, NA91o, NA92i, NA92l, NA92p, NA93r, NA93s, NA93q, NA94a, NA94c, NA94h, NA94g, NA95a, NA95e, NA95f, NA96f, NA96e, NA97e, NA97d, NA98g, NA98f, NA99d, NA00f, NAKH01e, NAKH02e, NK03d, NK04a, NK04c, NK04d, NK04e, NK05f, NK06c, NK06d, NK07e, NK07f]. U.S. [NK08e, NK09c, NK09e, NK09f, NK10e, Oom46a, Oom46b, Ora65, Osn77, Pai85f, Pan86a, Pfe82, Pol09a, PN09, Pos87, Pra04d, Pri76a, RL78a, Rep10, Rev92, RE68, Ric86b, Rin52, Riv96, Rob90b, Ros87b, Ros98, Rot97j, Rub08, Sci55a, Sau10, Sca88, Sch83, Sch54c, Sch98, Sch88b, Sco08b, Sea65, Shu87, Sim54d, Sim54e, Sin79b, Sol90, Bul03a, Bul08c, Bul09e, Bul09h, Ste85c, Str58, Sut96b, Sut68a, Swe78b, Tho53, Tod79, Tur52, Tur60, Utr87, Voh71b, Wal50, Wal99, Wri63b, Wri02, YG77a, You64b, You67b, vH87, Ano89e, Dow84a, Han98a, Mor79, Tod79]. U.S.-controlled [Ano84p]. U.S.-Soviet [Kin86]. U.S./USSR [Buc82]. U.S.A. [Bra95]. U.S.S.R [Ano84a, Dow84a, Zhe90b]. U.S.S.R. [Con88]. UAE [Mec16f]. überall [Pat79d]. Überterrorists [Ano99-28]. UCAM [Kal82c]. UCS [Kal82c]. UCS/UCAM [Kal82c]. Udaipur [Ano64p]. Udgaonkar [Moo93b]. UFO [Val73, Mar67a]. UFOs [Con69, Mun67a]. Ugly [Pra05b, Spa95a]. Uh [Ano96o, Aue02]. Uh-oh [Ano96o, Aue02]. Ukichiro [Smi54b]. Ukraine [Bud22, Doy22, Kas15, Kis93c, Kis93d, Kis94, Kun23, Les93e, Mar89d, Mar93b, Mar98a, Mye92, Rot97j, Str93]. Ukrainian [Sin14]. Ulam [Car03, Han92c, Wil76b]. Ulla [Bla84]. Ultimate [Bet76, Dal82, Dud75b, Har83, Ing52b, Neu07, Sal77, SD83, SG62b, Sai10, Gle94]. ultra [Len20]. ultra-targeted [Len20]. Ultraviolet [Kat54e]. Umbra [Bla02]. umbrella [Ark96g, FC86, Kan96]. Umbrellas [You78b]. Ummah [AH03a]. Un-American [Ano48u]. Un-Real [HWH78]. Un-word [Ano92t]. unacknowledged [GG16]. UNAEC [Ano49q, Cav48]. Unasked [CR99]. unassuming [Moo96f]. Unavailable [Bro81]. Unavoidable [Dee57]. unbalanced [BA89]. Unblocking [BV10]. unborn [BR79]. unburned [Ark97d]. UNCD [Jac65]. Uncertain [Cla74, Har60a, Kar99, Kop15, LK10, Pos71, Ric99, Sci49, OW15, Pos24, Ste83a, Tak11, H.84a]. Uncertainties [Ano78d, Neu58b, Swe93]. Uncertainty [Bre68, Cas92, Hib68, Man85, Swe02, Tho79c, Swe93, Sac84]. Unclassified [Adl85, Nat56, Rei63]. uncle [Pea88, Ano91v, Ano98n, Com74, Mas67]. Uncoiling [NA91i]. uncommon [Tsi87, Pol91b]. Unconquerable [Tay87]. Unconquered [Wei66d]. Unconscious [Ost63]. unconstitutional [McK89]. uncontrolled [Ste77c, Siv76]. Unconventional [Ack87h, Boy10, Osb83]. unconverted [Ano13e, Isa93c]. uncover [BWL21]. Uncovering [Ano14d]. uncovers [Alb89b]. UNCSTD [Mal79, Sei80, Wio79b]. undercut [Gil18]. Underdeveloped [A.51, Bar60, Bau60, Deu60, Mah59b, Ran55a]. Underdevelopment [Diw79]. Underestimating [Gre12]. Underfunding [Gor95f]. Underground [All59a, Ano03h, Bro60c, Dav79b, Ing79a, Rog71, Jeh60b, Kla19, PE24, Ste60c]. underlying [MS77c]. undermanaged [Dro22d]. undermanaged-consequences [Dro22d]. undermine [Bos88]. undermining [All14a]. underpinning [Hol77]. underrated [RW76]. Undersea [Cla16, Lac69, RC70, Gei16b]. Understand [Mur75, Hil16, Hil20]. understandably [Thi46]. Understanding [Cal84a, Con64, Ein48b, FLLW91c, Haw47, Hol65, IS15, Oba85, Pra02, Why71, Spr25, Gol55, Ker54]. understood [Rot92g]. Underwater [Pol01, Lau91b, Mer24, Sut16]. Underway [Foz62c]. underwear [Ano96n]. Undetectable [Ros59]. Undeveloped [Ded57]. Undone [Uhl03b]. Unearthing [Gao17]. uneasy [Fel79k]. Unequal [MK02b]. UNESCO [dF79a, Ano46-62, Ano52v, Bok48, Dav76b, Mal50, Noy47, Ric75b, Ano53-45, Ell71, Rab64f]. Unfair [Pik07]. unfeeling [Wes91]. Unfinished [Day75g, LM71, Smi84b]. unfolds [SCHD20]. Unforgettable [Joh81a]. Unfortunate [Kih47f, Rot00b]. unfought [Fre91c]. Unfreezing [Chu04]. Unfriendly [NA89s]. unfulfilled [Mül85]. Unhappy [Bar63, Hin69b]. Unheeded [Hal76a]. Unidentified [Hyn69, Sag67]. Unified [Kol07, Nor64]. Uniform [Bus46]. Unilateral [Asp87, Har84c, KH63, Lev63, Lon84, Mos89, Osg60, Sha61, Uhl01a, Wie84]. Unimportant [MY67]. Unindicted [Ark92]. Uninformed [Mar85a]. Uninspected [Lit78]. Unintended [CP93b, Dro22d]. unintended-and [Dro22d]. unintentional [SW21]. Union [Cra82c, Cra83d, Cra84d, Deu68, Dow83d, Etz07, Eva99, Fis83, Jor87, Klo85, Nye79, Ano50q, Ano58d, Ano84i, Ash50, Bal82, Bei82, Bru82, Cas85a, Cle82, Cra82a, Cra83c, Cra84e, Dal83, DK80, Eme71a, Fel79a, Git54, Glu50, Gol82c, Gra82a, Gra83, Gra86, Han91, HK02, Joh67b, Joh83b, Ken78, Kin86, Lap59c, Llu00a, Luc82, Pow82, Rab49e, Rab53d, Rau92, Ric52, Sem53, Tro58, Tsy83, Tur78, VA11, Wei83a, Yev86, vL86, Bei84, Ehr81, Mac62, Ses83, Som91, Soy89]. unique [Wei86b]. Unit [Wal85b, Kre89c]. Unite [Fel75b, Sto58, Goo85a, Hor89, Zhe90a]. United [Abb85, Ano52b, Ano57d, Ano78n, Ano78m, Dew55, Dow81d, Dow82a, Dow82e, Ekl80, Fel83e, Gol69b, Gre58, Kor92, Lal65c, Lal70, Mer90, Mes84b, Mos84, Nad79, Nie63, Per95, Rei63, Roc81, Sca91, Sla88, Smi54c, Sol68, Ste80a, Sto58, Uhl95b, Upt63, Van88, Wio84, Wur68, Ano88u, Mül90b, Al 11, AK98a, Ano46-105, Ano47c, Ano47y, Ano47z, Ano48-28, Ano48-29, Ano48-39, Ano51-28, Ano52-27, Ano54-27, Ano55r, Ano55y, Ano60e, Ano60c, Ano60d, AE85, Bar82a, Bec75, Bro47f, Brz12, Buc60, Car49b, Car49c, Car49d, Car49e, Car50a, Car50b, Car50c, Car50d, Car50e, Car50f, Cav56, Cor78, Cot19, Dal22, DRM^{+}09, Dje86, Eis54c, Eme71b, Eps82a, Eps82b, Eps92b, Eps97, Fre81, Fry57a, Gom83, Gru70a, Hes78, Hol84c, Hud76, Hud78, Hum58, Hun85b, KL65, Kal81]. United [Kar17, Ken83, Kih47c, Kih47d, Kih47f, Kih47e, Kih48a, Kih48b, Kih48c, Kih48d, Kih48e, Kih48f, Kih48g, Kih48h, Kih48i, Kih48j, Kih49, Kno60, KK19f, KK20d, KN16d, KN17d, KN18d, KK21e, KK21f, KK22e, KK22f, KKJK24d, KKJK24e, Lap51a, Llo71, Low85, Mau50, May17, Mec22c, Mec22b, Mia16, Mor50a, Mor54, Mor67, Mot83, Nit57, Noy46b, Osb49a, Pai82f, Pay70, Pea85, Pea91, Pea99, Rav83, Ros82, RE19, Sen57, Sal68, Sal73, Seg13, Sha87a, Sha78b, Shi57, Sig10, SL83, Smi21, Ste52, SSR83, Szi47a, TvH17, Tha00, Tra72c, Wei53b, Wei78c, Whi51b, Wie85, Wri61, Wu23, Wur92, Yua23, Zha19, dC78, vH16]. units [Ano88k]. Unity [Aug56b, Chu50, Deu67, Why59]. Universal [Ano51-27, Ano52x, Bro63b, Kat52c, Kre90, Ano91-41, Ano91p, Nov77b, Bok49]. Universality [Que52]. Universe [Lan58, Lan76, Lov69, Isa07, OW90b, Lib65, Smi58f, Smi58f, Sto59a]. Universities [Ass53, Ame53, Ano53-37, Ano68j, Boe57a, Boe57b, Don57, Fal68, Gla68, McC69, Mei53, Mil69a, Oro59, Shu78, Spr69, Wil68c, Ric74, And68]. University [Ack97, All51, Ano45c, Ano70g, Ano84d, Arc86, Bou02, Cam54, Doo64, DuB69b, Fed67b, Fed61, Hir72, Kah86, KAT^{+}71, Mar71a, McN62, Nor06, Sal70, SB91, Wel77, Wil71, Fri79, HFMS79, Lon86a, AAF^{+}52, Bla62, Che50, Gar88a, Hes78, MJ68, Pat80e, Sal73, Con67, Lyn69]. Unjust [Chi78]. Unknown [Eps75a, Tho80g]. Unknowns [Soc07]. Unless [Fel68b]. Unlikely [Poc05, TS99, GB11, Wur91a]. Unlimited [Gol73a, vA73, Brü55c]. unlinking [Sto78a]. unlisted [Ano98l]. Unlocking [BMPH18]. Unmaking [Hib06]. Unmasking [Tho12, Van80]. UNMOVIC [WS21]. Unnatural [Ack89h, HS96]. Unnecessary [Sai07, Ack91a]. Unofficial [Cra86b, Bat86]. Unpleasant [Dou69, Hel02]. unplug [Tay92]. unpopular [Tur77]. unpunished [Ano97i]. UNRA [Ano97m]. Unreal [Sel77]. Unreason [CSG71, Opp60, Rab71c, Smi72b]. Unreasonable [Pai04c]. Unreliable [Les94, Ano18, Gre18]. Unresolved [Bar80e, Swe79a]. unrestrained [Kis79b]. Unsafe [Pot97, Ano90-51]. unsavory [Jac82c]. UNSCOM [BWL21, Due21, Sel21, Tre21, Tud21, WL21, WS21, Wri99b]. unsecure [Den14]. Unseen [Ler80b, SAB^{+}16, Tay91]. Unsolvable [SG62a]. Unsolved [Rab54d]. unstable [Fis84]. Unsystematic [Gre55b]. untested [Wat77]. Unthinkable [Els90, Rot02f, Bur76, Nas83, Pea84, She85a, Bre84, SWC^{+}84, Fri62]. Unthinking [Els90]. Until [Fly01c, Isa94h, Nov77b, Ano78n]. untimely [KA88]. Untitled [Bea86a, Flo85]. Untold [Smi68]. Untrammeled [Col74]. untried [Moo92j]. Untying [Men92b]. unused [LB78a]. unusual [Mec16e]. Unveiling [Sto16]. unwelcome [Zit17]. Unwitting [Bar76b, Fin01, Gif13]. Unwritten [Gen84b, Haw73]. Upcoming [Ano94x, Ano94y, Ano94z, Ano94-27, Ano95y, Ano95z, Ano95-27, Ano95-28, Not15]. Update [Sme74, Wal90a, Ano82d, Ano95i, Ano95k, DM86, Lee97, NA87g, NA89m, NA98a, NA98b, Wal89c, Web93]. Updates [Ano01q, Ano01r]. Updating [HH87, Ros87a]. upgrade [Ger17]. Upholds [Ano48-34]. Upper [Hel70, Ano96l]. Ups [Cla74, Mar91b, NA89s]. Upset [NA91i]. Upsetting [She07]. upside [Lac17]. upstream [CYP14]. Upton [Ano46-68]. Upward [Isa00a, NA93j, RWK^{+}83]. Urals [Par80]. Uranium [Alb94g, Ano46-40, Ano50-42, Ano55s, Ano91-58, Bra50, Com08, Com75e, Dow82e, Fes77, Fis76, Gla76a, Gla02, Hec12, Hec20, Kat50a, Kat50b, Kat51a, Lan89b, Mar61b, Mar78c, Ner76, Nin59, NA93f, Rei68b, Rot55, Rot93a, Thi52, Tyl89, Wil80a, Zha15b, AS85, Ano92r, Buk93, Cal97, Cam79, Dic77, Die91, Fre77b, Hib95e, Mes97, Moo15, SD10b, Fra82, Bro55b]. Urban [Fri64, Gol61b, Haw51, Kra73, Lal66f, MM51, Owe79, Sin79a]. Urey [Coh81, Ure46c]. Urge [Atl47, Tol58]. Urged [Dys59]. urgent [MFN17, Mec17b]. Urges [Ano46-46, Ano47i, Ano48z, Ano50-38]. Urquidi [Bar67b]. Uruguay [OS81, Ote81]. Uruguayan [Bis14]. Ury [Bla84]. US/USSR [Ste82a]. USA [MNS46]. USAEC [Ano54t, Bos51, Sal53, Sim54d, Tur51]. Usborne [Ing56c]. USC [BBF^{+}82]. Use [ABB^{+}50, Ano46-107, Ano49-40, Ano53-46, Ano78d, Axe51g, Ber04a, Ber93b, Bie68, Cou52b, Cap98, CSF95, Cha58a, Com57a, Cou71, Fel67d, Fre68a, Fri64, Fry55, Got69, Har50b, Hir73, Hoe09, Isa85e, Jac73b, Kel64, Lac85, Len73, LD79, MF10, Nel94, Net75, Odh81, Rab47i, Sch47, Sch02a, Sei71b, Smi58d, Sti47, Sto76a, Thi52, Tod81, Uni57, Wal90a, WH72, Arc54, Ano79j, Ano91-36, Ark86e, Car94a, Car83c, Eps79, EP86, Fed76, Fel80e, Gro83, Hai81, Har79a, Hol20, HK21, KR12, Len21, Mil20, Mot83, Plo86, Pri80, Rav83, Sap84, SSR83, SCHD20, TBBS85, Tin23, TC83, Wei83b, Bun82, ZG83, Ano83k]. Used [Com46, Luc90a, Gau24, Mar82c, Moo15, RL14]. Useful [Han96b, Hud78, Moo93g, Tho90]. Useless [Mey61]. User [Aeb48, Ano46r, Gav51]. Users [Rei50]. Uses [Ano46-58, Ano48-31, Ano54u, Ano56j, Bha55, Bru50, Coc59, DR71, Foz62b, Fra57, Hur78a, IS67, Pre62, Rho46, Ste60a, Tel51a, GP18, HK21, LaF96]. Using [Isa86f, Hic22]. USSR [Bus85, McC93, Rab64a, Ano78i, BS57, Buc82, Cha78a, Coh72, Foz62d, Khr61, Kra59, Mul48, NAC87, Rab52h, Ste82a, Tur60, You67b]. usurps [Bro87c]. Utilities [Lee74, Par46a]. Utilization [Lon52b, Rub74]. Utopias [Rus62b]. Utter [Cap98]. Uwe [Cra84b]. Uzbekistan [Ano94-28].

V [All56l, All59l, Bru57, Ekl80, Fel73a, Hol86a, Leb91b, Lyn69, Pla57a, Rab63a, Rab67a, Sin81, Van88, Ano90r, Ano92x, Ano94t, Ano94-28, Ano97n, Ano98k, Gra94, Khr94, Lan90l, Men91a, Mil02, Nad98a, Per92a, Rot02e, Bul01c, Ano51-29, Bro86h, Dje72a, Ell56, Fin56, Gan72, Rab52i, Fis55]. V-2 [Fis55]. VA [Fre93b]. Vaclav [Mye93b]. vacuum [Isa89e]. vadis [PG84]. Vail [Ano81b]. Val [Ano53v, Lap54a]. Valedictory [Fel84h]. Valerie [Lyt79, Mec16b]. Valid [Fas66]. Valindaba [AZ89a]. Valley [Dro22e, Dro24d, Hec22, Kam14, Roc78, RHKK78]. Value [BMPH18, Bur57, GLMS60, HB63, Led71b, Man57, Opp51a, Pla73, Sea63b, Ano93y, Gel90, Ham76, Oel14]. Values [Ano96o, Bar63, Ber77b, Ehr82b, Ell47, Fre67, Jor67, KK72, Opp59b, Sag86, Spe72, Spe66, Wei66c, Gou24, McD74b, Zac76, Haw58a, Saw81b]. Values-free [Ano96o]. Vampire [Rot92t]. Vance [Her83a]. Vancouver [Ano80n]. vanden [Ano90o]. Vandenberg [Ano48-37, Van52]. Vanguard [AC02, NA92e]. Vanished [Wil68b]. Vanishing [Cha68b, Gom81c]. Vannevar [Ano98g, Cha84, Lak66, PMRP50, Ano85u, Sha84d, Zac92]. Vanunu [BED93, Day92, Day96, FP89, FJ91, Mos88, Rot98g]. Vardø [Sel00]. Variable [Wor69]. Variations [Soh62]. Vary [Ano64q]. Vasil’evich [Hol86a]. vast [Fre77b, Fre91d]. Vatican [Jac74b]. vax [Fie24]. VDTs [Har91a]. vector [Sch80b]. Vehicles [Lal65g, Net70, Sut61, Spe18]. Veil [Dav79b, Har88c]. Vela [Pre63]. Velikhov [Bou87]. Velikovsky [Kol75, Mar64i, dF79f]. Vendreyès [Ano85h]. vengeance [Tur77]. Venice [Ano80c, Ano80d]. Ventilator [Cro18]. Venture [Coh53a, Fuc70, Dro24d]. ventures [Rot89b, Sch88b]. Venturesome [Gre67b]. Venus [Foz63a, Lew67b]. Vera [Rab72a]. Verbal [SNB^{+}54]. verdict [Gor96f, Opp54a]. Verdon [Say84]. Verdon-Roe [Say84]. Verge [Gar81]. Verification [Ano63-28, Ano86k, Dre91, FM01, Flo90, Gol88e, Hin18, Lal64d, Ric98b, Sar78, Sco70, Wri02, Ano87-44, Buc83, DeV92, Ger13, Gol85a, Haf85, Hec86, Hic22, Ibr13, Kre85b, Kre87a, Lei83a, Pai83e, Pif18, Rot92a, Sam18, Tim87, Tuc10b, Wil83, WL21, Zer13, Ano86k, Ano86k]. verifies [Roc18]. Verify [Mei01, Dor90, Hei89, Hic22, JH16]. Verifying [Cha75a, KKS22]. Verlag [Rot97c]. Vermont [McC81]. verse [Isa89a]. Version [Ano48-33, Ano79i, Ano79m, Ano79f, Ano79g, Ano88g, KS93, Tsi89a]. Versus [Cav47, Ano84a, Dow84a, Fre88, Har54, HL49, Nat85a, Pat81a, Ric53, Spe54]. Vertical [Yor82]. Very [Ano97k, Hud70, Sch99, BV98, Dro23k]. veteran [Ano80h, DD79, Tho79b]. veterans [Cam88, Kal83b, Mar83c, McK83]. Veto [Ano47o, Ano68c, Ing46, Sen47, Shi47b, Ros90a]. Vetoes [Wri02]. vets [Man95, Rot92n]. Vexed [Dav48]. VI [Ric52, Fri56, How56, Mey56]. via [BMPH18, HC87, JH16]. Viable [Rab69f, Fel85c, Pan98a, Rob18]. vice [Gut93a, Ano91-33]. vice-king [Gut93a]. Vicious [Bol63]. vicissitudes [Arb18]. Victim [BED93, Loe02, Osn77]. Victimless [Bar81g]. Victims [Ano46-67, Abr86a, Ste90a, Zhe90a]. Victor [Adl69, LL90, Bur48]. Victorian [Bri80b]. Victory [Bou02, Ger67, Haa77, Rab46g, Ack95, Ano93s, Isa85b, Isa91o, Moo92e, Nel92b, Off02]. Video [Hal93a, Ano91-41]. videoski [Ano90-30]. vie [Bon57, Bon58a, Bon58b, Har87d]. Vienna [Ano58l, Ano58k, Ano59d, Ano58l, CW81, Coo85, Kap79, Lea89, Lew68b, Lew70d, Ran61, Wio79b]. Vietnam [Ack95, La 91, Sol68, dF79d, Ano67f, Ano67o, Ano80h, Bar65, BS67, Bro66, Bur74, BTCSC^{+}66, DD79, Fed67a, Fas66, Fot91a, Fra65, G.65, Gol70c, HT03, Ing67, Isa05b, Ist67, KL65, LW69, Mar65a, Mar65h, Mar65c, Mes71, Mor65, Rab65g, Rab72c, Rab73a, Rus70b, Sen67, Tho79b, Whi64, Gom71, Wur68]. Vietnams [Gol69b]. View [Ass63, Ano52g, Arb82, Aro60, Aue01, Ben69, Ber64, Bin10, Bro72, Buc60, Bur65, Cah71, Cal66, Coh74, Col59, Dro16b, Exe57, Edg52, Far71, Fin75, Fit75, Gad72, Gil69b, Gil69a, Gof71, Gom65, Gom75b, Gre67c, Hed65, Het55, Hey72, Hul78, Jam58, Kis93c, Kno60, Lal70, Lei72a, Lev63, Lew49, Lew71f, LM71, Lop66, Lov73, Nar67, O’B69a, Pag69, Pas55a, Pra06c, Rab50a, Sch00c, Sen67, Sim53, Smi72a, Spu73, Tod64a, Yor70, You64a, Zac50, Ano84i, Ano87i, Boy86a, Car83c, Col95, Coo78, Fit11, Flo76b, Fra47, Goo85b, Gro83, Hec86, Her78b, Hew81, Hop77, Joh96c, Kap85, Kap83b, Kla19, Kok88b, LaF90a, Les93e, Mar94, May85, Rav83, RKK22, Röl81, San76b, Sap84, Sch00d, Sha92c, Smi85b, Spr19, SSR83, Teb85, TC83, WW60, Zam96c, Sel63]. Viewpoint [Ano49-37, Swa49]. Views [Ano46y, Ano46-31, Ano49-30, Ano71f, Ber49, Ber72b, Bet69b, Bet77b, Dys69d, EBU^{+}52, Fel72a, Fin55, Foz63b, Ful69, Har72a, Har47, Har64b, Hum72, Lan61b, Lap72, Mar53c, Ore59, Pla65, Pol49, Rab59m, Rat69, Ros59, Sco72, Ste69f, Sti61, Tay69, Ure46a, WC47, Yor69, Zei53, Alb87a, Avi83, Cle87a, Fal87, Gou75a, HC87, Hui87, RL78b, SH95, Sha87d, Sub87, Tre19, Väy87, Mos81b]. Vigil [Ano81i]. vigilant [Ano92s]. VII [Ano51d, Ano62u, Bel57, Gan52, Mor57b, Sin56, Vis56]. VIII [Ano52e, Adl57, Eps52, Fra57, Pol57b]. Viking [Wil71, Jac71a]. Vikram [Sea72a]. village [Pal77a]. villages [Usm79]. villain [BED93, Dro22i]. Vincent [Kat53a, Cop65]. Vindex [NA93j]. violate [Ano94s, Ros94]. Violating [Gof93c]. Violations [Fis62b, Car94d, Duf86, Gol84a, You85a]. Violence [Day75h, Mar64j, Mau77, Pre75, Pre76, Rab69j, Sch67b, Tam74, Whi77, Ano94n, Eva75, Lif13, Tho80i, YWS20, Dev69]. violent [Ano06q, Dav77a]. violins [Ple91]. Viorst [DeB81b]. virtual [Lin12]. Virtues [Wea75]. Virus [Rab68g, Sue68]. Viruses [Ang60, Bea59]. Visa [Ano53-46, Ano54h, Ano55o, Ano55x, Aro52, EBU^{+}52, Fed55, Gil55, Kal87, KASN52, Wei52a, Wei52b, Doy53]. Visas [Ano54-27, Rin52, Wei54b]. vise [Pav98a]. Visible [dF79e, Gal13]. Vision [Dav74, Bul09a, Uhl01b, Wil73b, You73, Ano96m, Arn80, Jac83, Led94, Sho92b]. Visionaries [Moo95k]. visionary [Bes85]. Visions [Ack95]. Visit [Fre63, Lon52c, Oli66, Sim46b, Fel82b, Mac77b, War95]. Visiting [Ano85a]. visitor [Pal77a]. Visits [Axe60]. Visual [Why59]. Vital [Nai81, PvH74, Kla95]. Vitals [Ano96-29]. Vitrification [Rot92u]. Vivian [Sti83]. Vivienne [Say84]. Vivre} [Hul76]. Vladimir [Gol93d, Dro25c]. Vladivostok [Lon75]. vocabulary [HNR12]. Voice [Ano73d, Ano90-30, Aue06b, BBH^{+}73, Fel73d, Hut61, Ing74, Isa89e, Lap73b, Mei73, dSP60, Ano81m, But98, Gol89, Mos79, Ram13a]. Voices [Gre67e, Sim84b, Squ77, Ind81, Ano90o, Gen72]. Voicing [Sol76]. Vol [Ada52]. Vol. [Cra84b]. volcano [Dro22h, Ded60]. volcanology [Ano92s]. Volkmar [Wri63a]. Volsky [Gut93c]. Voltage [Sch75a, YY74]. Volume [A.54, And65, Ano49j, Ano50g, Ano51d, Ano84g, Ano87r, Ano87n, Ano88d, Cal84a, Gre67b, How65b, Kat54c, Nie63, Shu78, Sch04c, Ano52e, Ano53j, Ano54f, Ano55d, Ano56e, Ano57e, Ano58b, Ano59b, Ano60g, Ano61e, Ano62m, Ano64h, Ano65d, Ano66a, Ano67c, Ano68b, Ano69b, Ano71a, Ano72d, Ano73a, Ano74a, Ano75b, Ano76a, Ano77k, Ano78h, Ano79b, Ano80e, Ano81d, Ano82b, Ano83e, Ano84c, Ano85e, Ano86m, Ano87t, Ano88f, Ano89h, Ano90q, Ano91t, Ano92k, Ano93o, Ano94i, Ano95h, Ano96h, Ano97h, Ano98i, CAH84, C^{+}87a, C^{+}87b, C^{+}89, NBF94, For80]. Volumes [Ano78o]. Voluntary [Sta62a]. volunteer [Gau24, Jan81b, MJ80]. Volunteers [Ano99-30]. Vonnegut [Bri86]. Voslensky [Lon85c]. Voss [Ber77b]. Vote [Isa02b, Isa03d, Ano96-30, Isa88c, Kor89]. Voter [Bre85]. Voters [Ano46-50, Mes84a]. votes [Ano89i, Pik96]. voting [Cal24]. Vox [Gom62b, Gom62c]. Voyage [Lew72a, Mar58a]. Voyages [Fol75]. Vremia [Ano92x]. vs [Abr73, Ada73, Alf81, Ano72g, Ano87-31, Bet69b, Bou98b, Cha88c, Com75g, Cus49, Dow77, Dre46, Hig48a, Hym71, Jac73b, Kre77, Lal66d, Lev12, Lev53, Man84c, Nor74, Por82, Rab70b, Rab72c, RD75, Ric79, Ros75, Ros87b, SC75, Sta87, Ste46, Szi64, Thi58, Wal77, Wal76]. Vu [Wei04, Ano94k]. Vulnerability [Coa47, Koc77, Ano86-28, Mor84b, Rot88, Wol17]. Vulnerable [Ano99b, Bro57d, Ric12]. Vyacheslav [Leb91b].

W [All56l, Ano50-41, Ano78n, Ano78l, Ano81c, Ano87n, Bal79c, Ber73c, Bro47c, Bru57, Cal62, Dia84, Fel83d, Fer57a, For80, Han94, Has57, Jas81, Kap67, Lak66, Lap55a, Mar54e, Mil85c, Nor64, Pom67, Rad83, Rei78, Rep86, Rot04d, Sch71a, Sch81a, Scl79, Sim87a, Smi58e, Tod79, Tsi72a, Tuc97, Win81a, Wol55, Wri63a, Wur68, dF79a, SS05]. W. [Mar82b, Pan00, Rab60h]. Wackersdorf [Ano89l]. Wadsack [Len21]. Wadsworth [Alc63]. WAES [Sch78a]. waffle [Ken85]. wage [Lop90]. wages [Rot93s]. Wagg [A.52]. Waging [Ark85i, Han87a, Pre92, Sol91]. Wagner [Fin90]. Wagon [Isa03f]. Wainhouse [Pom67]. Wait [Ura01, FTvHW76a, Smi93d]. Wait-and-See [Ura01]. Waiting [Lyf61]. walk [NPS90]. Walked [Pra03b]. Walker [Cap98, Moo02d, Rot93a, Mes82]. Walking [Fel82j, Gey94, Gur88a, Lar99, Rab56m]. Wall [Kau82, LaF90b, Lyn69]. Wallace [Ano52q, Dre46, Kre89c, Wal46b, Wyl56]. walls [Isa98b]. Walt [Moo96b]. Walter [Ano78n, Bri86, Dia84, Fis55, Hol86b, Kat53a, Kru74, Mil85a, Rei63, Rus51a, Smi85d, Swe86, You66c, Ano52-28, Ano87-27, Jac73g, Moo95l]. Waltz [Lew70d, Isa93h]. Walzer [Chi78]. Wandering [Gar91]. wand’ring [Gor96b]. Wanes [Ano62-28]. Wang [Dro22a]. Want [Cha63a, Fis62a, Hit03, Kal77b, Man66b, Mic68, Pie68, vW58, Ano91-37, Ano94m, Ano98m, Bar81e, Pet96, SOE21]. Wanted [Ano79o, Ano96q, Ark02c, Mor53b, McK25b, Moo96e]. Wants [Hay93, Min94, Smi94b, Ano91v, Ano95i, Jon91a]. War [Ahm03a, All68, Ano46-91, Ano49-35, Ano68h, Ano86i, Ano87s, Ano89c, Ano89g, Ano90m, Ano91-35, Ano91q, Ark85f, Ark98d, Ark00d, Ark02c, Aro56, Bal50, Bal56, Bar80i, Bar78c, Ber76b, BD79, Bla84, Bou02, BB79, Bri87, BEI^{+}55, Bus85, BBHGA57, Cha06b, Coh60, Com57a, Coo71, Cra83b, Cro65b, Dav57a, Dow81f, Dow81d, Dow82f, Dow83d, Dow84b, Ede89, Eds62, Ein79a, Eis68, Elm68, Fed67a, Fel76a, Fel77o, Fly99c, FW64, G.65, Gai92, Gar97a, Gla63, Gol93a, Gol88a, Gol69b, Gra62, Gra60, Gra93, Gus83, H.84b, HP80, Hal91, Han92b, Har04, Hig50, Hod00, Hos87a, Hos87b, Hut46a, Ing62, Isa88e, Isa02f, Isa03e, Isa03g, Joh64c, Joh83a, Joh97a, Joh99a, Jou52, Kal82c, Kar80, KS58, Ken79, Kin83, Klo85, Kno55a, Koc57, Kre89c, Kre98, Kri85, LaF90a, LaF62, Lap56d]. War [Lap59i, Leg58, Lei74, Lei68, Ler80a, Lon78, Lon99, Lor57, Mac51, Mar64j, Mar58b, Mar61d, Mar89e, Mar98c, Mar02c, Mat57, McC56, McG67, McK99, McN63b, Men64, Mic62, Mil85b, Moo99f, Mor56, Mor85a, Mul65, Ols62b, Olu71, Opp68, Ost63, Phy80, Pal77b, Pet50, Phe61, Pra03a, Pri86b, Pri89, Rab51j, Rab58a, Rab71a, Rad92b, Roc81, Ros90a, Rot58a, Rot91a, Rot04h, Rus46, Rus58b, RE63, Rya89, S.55b, Say84, Sch60a, Sha84b, She89, She61, Shi54c, Sim84a, Sim87a, Smi83, Spe55, Spi74, Swe93, Szi62a, Tel46a, Thi58, Tho80g, Tic94, Til64, Top58a, Tre07, Uhl95b, Uhl98a, Van03, Van88, Was64, Woo01d, Wri47b, You82, Zha99a, Abr80, Abr84, Ack91a, Ack91e, All14a, All17, Ano90-44, Ano92y, Ano93x, Ano93-27, Ano94w, Ano97i, Ano97n, Ano15a, Ark86f]. war [Bar77c, Bar80h, Ber81, BM99, Bru82, Brz87, Bud22, But84, Chi17, Cho81, Cla77, Col79c, Cor95, Cra82d, Cum98, Dal17, Dav82, Day77i, Dow82h, Doy22, Eav16, E^{+}84, Ehr84, Eme76, Fal79, Fel79a, Fel81l, Fis81, Fly98a, Fra79, Fre91c, Fri82, FWCL20, GLH^{+}79, GvH23, Gra89, Hay79, HPR91, Isa90h, Iwa82, Jac82a, JM85, Jon13, Kal83c, Kal89, Kin84b, Kis79b, KR12, Kre86c, Kun23, La 79, LR79, LL83, LL12, Lif80, Lif13, Lop90, LJ17, MDB81, MK02a, Mar82c, Mar88, Mea93, Mec17c, MW91, Mil79, Mil83a, Mil85d, Mit83, Moo93a, Mul80, NR85, Nin88, NA89m, Pon82, Pug76, Pat80d, Pei82, Pod16, Pol80, Por82, Pow93, Pre92, R.83, Rin84c, RT12, RTB^{+}19, Rus20, Sal79a, Sas17, Sha95a, Shi24, Shu81, SW21, Sok22, Sol91, Som84d, Ste81a, Sto16]. war [Thu79, Uri94a, Wal90e, Wei13, Wei82b, Wes81, Woo86, Zel92, Zuc82, Zwe85, vH^{+}90b, vL47, Wal81b, Ada52, Ano47-28, Ano52g, Ano58i, Ano90i, Ano91j, Ano91k, Ano92t, Ano92-33, Ano97m, Arb77, Ark93b, Arn91, Bra95, Bri86, Cra86b, Cum92b, Cum97b, Dal22, Deu83b, Dia83, Dow81f, Dow82f, Dow82g, Dow82h, Dow83d, E^{+}84, Fis62a, FLLW91a, Fot91b, FHN^{+}46, FHN^{+}63, Gol70c, Hal93a, Har86a, Hol96a, Isa90-27, Isa94e, Jac82b, Kal88, Kal95a, Kle91, Lop91, Mar89a, McD85, Moo92e, Nat55, Nel57, NK16, Off02, Pai04a, Pat87, Pei53, Pra92, Rab50b, Roc81, Ros81, Rub08, Sch19, Sei50, Sha91b, She61, SW21, Tar23, TS22, Tsi89c, Uhl97b, Whe90, Wie67, Wri63a, Wu22, Ack95, All54b, Ano87p, Bar58a, Bar79b, Bou98b, Boy87, Boy89b, Boy91, Coo81, Dia83].

War [Dow86, Dow81f, Dow81c, Dow81h, Dow82b, Dow83b, Dow84b, Dug91, Eva86, Fer69a, Hal93a, Ise92, Kra94, Lac85, Lev84b, Mil86b, Moo96b, Moo97a, Mor82a, Pan63, Per91a, Rab98, Rod85, Rom82, Sha59, Sha98b, Smi85d, Ste89e, Sto59b, Uhl95a, Wal85a, Zag91a, Bel93, Bet65, Bri89, Dow77, Dow79e, Lep62, Tho62, dF79g]. War-fighting [Bar80i]. War/Education [Kal82c]. War/Peace [Dow84b, Kal83c]. Ward [Mei77a, Nas62, Sto55b, Ven80, Red69]. Warfare [Ame47, Ano52c, Ano53-46, Ano87q, Ash60, Bar61b, CT78, Chi59a, Chi59b, Cro65b, Dub60, Fed64, Fer57c, For49, Fre95, Gla60, Ims60, Kre89c, Kri85, McE88, Mer46a, Mes70, Mes72, Phi54, Rab59i, Rab60h, RKR55, RK55, Rid50b, Rus70b, Sch69a, Sei69, Sel59, Tel57b, Thi47, Ano68l, Ano88h, Ban11, Bar76c, Ber85, Chi17, Han87a, Lin19, Mie77, vCM89a, vCM89b, Per91a, Por19, Sch18, WS83, YSW21, Sid69, WSH^{+}84, Kel98a, Tuc97]. warfighter [FG14]. Wargames [Gol88a]. Warhead [NA88o, NA91b, vH91, Ack88m, C^{+}87a, C^{+}87b, CN90, Hic22, NA87g, PC92a, Sco84a]. Warheads [AZ90, NK04a, NK09f, Sch05e, Sho89, Ack88n, AP88, AT88, Ark86d, Buk97c, CAN88, Hor88b, NA88a, NA92k, NA93i, NK15, Ram88a, RS88c, Ste88, Sut91, Wag88, vH16]. Warlords [Kil78]. Warm [Gus69, Isa92i, Lon52b]. warmer [Dro20b]. Warming [Kel78, Lan90i, Sin92, Smi90a, Ano90-34, Ano91-36, Ano13a, Kel20, Leg92b, LG15, MM90, Nor11, Nuc20, Sci92, Sch92a, Sto17c, Wea11]. Warmongers [Kih47f]. Warned [Bul01a, Coo76a]. Warner [Pat79a, Sco74, Wol63]. Warning [Bro56a, BFM^{+}49, Gai69, Smi53, BØ79, Day76b, Eld92, Fel82k, McA84, Mye93b, Qui84, Roy76, Reh14, Sin92, YWS20, Sti83]. warnings [Dro23h]. Warnow [For80]. warpath [KA91, NA87d]. Warpaths [Van91]. Warren [Gru70b, Has57]. Warring [Smo89]. warrior [Osb83, Rot79b, Ano13a]. Warriors [Hal93a, Hit07, All18, Cla91, Fel80c, Fel93, Gor96b, Rot91a, Pre86]. Wars [Bar80m, Che88, Chi78, Deu83b, Gol88g, Han92b, Har88c, Kla90, La 91, Sar85, She05a, Cha88a, Eké21, Fra89, Jon91a, Lan90b, LP92, Møl90, Rot95i, Sch81c, Wil86a, Mec19h, ATL85, Aft86, Ano86i, Ano97k, Arb85, Ark86g, Bal86a, Bro85i, Bro87a, BS85, Fel85c, Gla85, Gom87, GK94, Gur88a, Gur88c, Har86d, Hec19, Her87, Isa88c, Isa93g, Isa95d, Isa96f, Jon93a, Ken85, KW86, Kub88, Lac85, Lan90a, Lea87, Led85, Men96b, Mon89, Mye86, NA89j, Ric86b, RK87b, Rot92o, Rot94k, She85b, She86a, Smi87b, Smi87c, Smi00, Swe87a, Swe87b, Tir86, Win87, You87b, vH85, Rot97b, Sha95b]. Warsaw [Fis83, Sta85, Ano82t, Ano87-31, Ano96k, Hay81, Lal64e, Sha85b]. Warski [NA87f]. Wartime [Man74, Str05f, STS^{+}49, dH75, vL48a, vL48c, Sha87c]. Was [Dud75a, HPR91, Kap74, Leb88b, Rin75, Rot96b, She87, Spi57b, Ano91-63, Ber01, Cir05b, Coh98, DM24b, FE23, Hib96a, Hir89b, HM90b, Hol96b, KAS96, Kno90, Loc89b, Rot98f, Sch76a, Sic96c, Wee95, Wio79b, Wit95a, Wri89, ZRG83]. Washington [Dow80h, Ato50, Ada86c, Ark83c, Ark83a, Ark83b, Ark84a, Ark84b, Ark84c, Ark84d, Ark85b, Ark85c, Ark85h, Ark85d, Ark85f, Ark85e, Ark85g, Ark85i, Ark86c, Ark86g, Ark86d, Ark86a, Ark86b, Ark86i, Ark86e, Ark86f, Ark86h, Ark87c, Ark87d, Ark87e, Ark87f, Ark87g, Ark87h, Ark88b, Ark89d, Cai89, Eid66, Flo88, Flo89a, Flo89b, Gar84b, Gro53d, Gro54a, Gro54b, Gro54c, Gro54d, Hay81, Hin69a, Hin69b, Isa78, Isa81b, Isa81a, Isa82c, Isa83a, Isa84b, Isa85b, Isa85c, Isa85d, Isa85e, Isa85f, Isa86a, Isa86b, Isa86c, Isa86d, Isa86e, Isa86f, Isa87c, Isa87f, Isa87e, Isa87d, Isa88a, Isa88b, Isa88d, Isa88c, Isa88f, Isa88e, Isa89i, Isa89j, Isa89k, Isa89l, Isa89m, Isa90-27, Isa97f, Khr10, Kin78a, Kin81a, Kin81b, Kor17, Kre86b, Kre86c, Kre86e, Kre86d, Kre87b, Kre87e, Kre87c, Kre87d, Kre87f, Kre88d, Kre88e, Kre88f]. Washington [Lan79, Lev84a, Lew72d, Lew80a, Lew80b, LN83, Lin84b, Lon87, Mag88b, Mar64g, Mar64e, Mar64h, Mar64f, Mar64i, Mar64b, Mar65a, Mar65f, Mar65g, Mar65h, Mar65d, Mar65b, Mar65c, Mar66b, Mar49f, Mar49c, Mar49d, Mar49e, Mar51, Mar53a, Mar53b, Mat79, NCA88b, Pai82b, Pai83f, Pai85f, Pai85c, Pai85d, Pai85b, Pif18, Pri69, Ros88, Spe87b, Spi58, Swe78a, Swe79a, Swe79b, Swe80, SM89b, Tam71b, Voh71b, Voh71c, Voh71d, WB87, Wal87b, Wei78a, Zam96c]. Waskow [Wol62]. Wasn’t [Ano95n, Ark89c]. Waste [Abb79, Alv05, Ano71g, Ano91-37, Ano94m, Dow80j, DSZ74, Fla60, Gol11, Gun04, Ial18, Kee76, Kra73, Mas76, Maw77, McC09, Met83, Mic74b, Pai04e, Bul08q, Wol58b, Zin79, dL59, Ass80, Ack94, Ano90-41, Ano93s, Bet98, BM22, Car94c, CSF95, Car97, Col79b, EG24, Gib91a, GMV^{+}24, JS79b, Joh80a, Kla19, KM18, Law84a, Law85, Mac11, RS88c, Rom92, RG13, Rot95l, Sal90a, Shk92, SF94, Swe79a, TPPB21, Wal21a, Wil93, vD78, vHW76b, vHW76a, KM18, Met83, Law88, Wal84]. Wastebasket [Alb61]. Wasted [GW02]. Wasteful [Fei07, Mec19e]. Wasteland [Rod70]. Wastelands [Pre96]. Wastes [Ano78n, BKSG60, Spe74, Ter58, WZSA73, Woo73a, ZSA73d, Hun84]. Wasting [Bra17, NA90t, Utl51, You77b]. Watch [Ano90-50, Ark99d, Isa01c, Lor99, Moo01d, Ano91m, Ano91-33, HIKWB10, Kal82f, NA87e, NA87f, NA87d, NA87g, NA87h, NA87i, NA88c, NA88f, NA88g, NA88h, NA88l, NA88m, NA88a, NA89b, NA89c, NA89j, NA89a, NA89o, NA89s, NA90c, NA90d, NA90j, NA90m, NA90q, NA91b, NA91i, NA91m, NA92j, NA93j, Dro22a]. watchdog [Flo76b]. Watchdogs [Lad81]. Watching [Ded78, FP05, Ric06, Ros06a, ZvH01, Leo19, Rot92c]. watchmen [Ded78]. Water [Ale69, Ano67i, Ano69h, Ano88c, Bet46, Bet75b, Car84a, CJ01, Dos72, Eme65, Fox61, Gat72, Gle21, Gol09, Isa93l, Kne65, Lee74, Lew65b, Mye89, Mye02, Nac65, Nac69, Nac72, Shi24, Smi69a, Wol65, AH22, Bet20, DJ13, Dro24g, Gle24, IS15, Ner80, SJ82, Ste91b, Dow84d, Tho84a, Haf79, LD79]. Watergate [Day75k]. Waterloo [Isa04b]. Waterpower [Dow83c]. Waters [Hol05a, Val97b, Isa92e, Lam87]. watershed [Hib12, Bir87]. Watkins [Ano90-63, Ano91-38, Ano91-64, Lan90k]. Watson [Rab68a]. Wattenberg [Dow81g]. wave [Ano95n, NA88o, Pai82a]. Waves [Chu03, Han92c]. Way [Amb01, Ant74, Cha62, DS49, Eba72, Fel83d, Fly05b, Gil69b, Hol06, Hur75, Hut53b, JvH09, Kat92, Kor75, Lev70, MvHKN01, O’B69a, Ols59, Rab46f, Rab48i, Rab53e, Ros75, Sie05d, Sie06f, Sto56a, TV71, Abr79, AT88, Ano93w, Ano97j, Ano98k, Ano14f, Ano14g, Arb19, Bos88, Dro22b, Fel77j, Hol96b, Joh97b, Kap81b, KAS96, KR12, Lat79, Mar77, MvE13, NA90k, Sep23, Str94, Zag93c, Dow80e, McC54b]. Waymack [Ano49m]. Ways [Bak75, Dra69, Sau01, Hei11, Leb85, MZ96, Plo86, Ric98b, Tur18, Ung83, Ano83b]. wdie [O’K84]. weak [Rot92s]. weaker [Ano86f]. weakness [Dav77a]. Wealth [Gra76b, McK74, Pon67a, Kis93b]. Weapon [Ano49f, Ano87r, Ano88d, Ano88e, Com57b, CK20, Mit70, Moo93b, Neu07, NA95a, Old62, PI92, Rab57d, Riv62, Bul09a, Wal08, Ano80i, Ano86-34, Ano95l, Fel80k, Fra89, HCH16, Isa89m, Kid91, KN14d, LB78a, LC94, ML82, Pan91, Pai80a, RP82, She94, Vog90b, Wal90e, vH91, Ano78p, Her82a]. weapon-free [RP82]. Weaponeers [Pai04e, Pri81a, Moo96e, Wei94d]. weaponization [JFB19, Gho15b]. Weaponry [Bur90c, Cle70, Dow81h, JL74, Ros90b, Bou87]. Weapons [Ato54b, Abr04, Ale60, All57k, Ano50-29, Ano56i, Ano57k, Ano59i, Ano60d, Ano60k, Ano65g, Ano66r, Ano75i, Ano82d, Ano83g, Ano84g, Ano85d, Ano85c, Ano86-36, Ano87q, Ano90-64, Ano91s, Ano91-66, Ano93l, Ano96v, Ano96w, Ano98h, Ano10a, ADK91, AK01, Bal86b, Ben01b, Ben78, Ber04a, Bor53, BHH11, BES02, Bro57b, Bro57c, But97b, Cou52b, Com51, Cal84a, Cas02, Cav57, Cla54, Coh62, Dah78, Das85, Dav60, DM85, Dow81a, Dum80, Dys60, Els90, Fed67a, Fed59b, Fel64, Fra66, Fry55, Gal56, GSB^{+}83, GP84, Gla02, Gol70a, Gra84b, Hal75b, Han94, Her85, Hir50, Hua02, Hum60, Int50, Ing52b, Isa03f, Kal64, Kal82b, KS58, Keu83, Kra99, Los57, LM80, Lan90o, Lap55a, Lap59b, Lap61b, Lap70, Len73, Len11, Lie62b, Lym96, M.85, Mar78a, McL11, McM51, Mes64a, Moo95c]. Weapons [Moo97a, Moo98b, Moo01c, Moo01f, Mot49, Myr75, Neu58b, Nor55, NA87e, NA87f, NA87d, NA87g, NA87h, NA87i, NA88c, NA88f, NA88g, NA88h, NA88l, NA88m, NA88a, NA89c, NA89j, NA89a, NA89l, NA89o, NA90c, NA90m, NA90q, NA91b, NA91i, NA91k, NA91m, NA93j, NA93q, NA96e, NK04e, NK09b, Opp53, Opp63a, Ore58b, OS00, Par86, Pau62, Per92a, Pin57, Pre96, Rab50b, Rab59n, Rez07c, Rob75b, Roc81, Rot58a, Rub08, Sea89b, Sei53, She89, Sil66, SIF66, Sim54d, Sim54e, Sim11, Smi77b, Smi92d, SS90b, Sto60, Tam74, Tay56, Tel47c, Tho84c, Tuc01, Tur53, Ull63, Van03, Veg97, Vel84, WS91, Wat70, Wel77, Wig58, Wig62b, Yon68, You65a, ZvH01, Abr87, Ach12, Ack88n, Ada82, AT92, Agn77b, AZ89b, Alm85, AS85, Alv14, ASB^{+}84, And85, Ano82m]. weapons [Ano83c, Ano86g, Ano87k, Ano87y, Ano87-45, Ano88q, Ano89u, Ano90-38, Ano90-39, Ano94s, Ano14d, Arc86, Ark83b, Ark89c, Ark96g, Bac81, Bai80, Bar80i, Bar92a, Ber87c, Bev86, BWL21, Bur90b, Car13, Cha16, Cha85b, CAH84, C^{+}87a, C^{+}87b, CAN88, C^{+}89, Cou77, Day77c, Day77k, Den14, Dom82, Dow82i, DG07, Dro16a, Dro16b, DW90b, Edw12, Eps79, Fei81b, Fei90, FvHA86, Fel76h, Fel78k, Fel79n, Fie87, Fie91, Fis91, Fit11, Fra21, Gao17, Gar78b, GR79, Gar88d, Gar92, GN10, Gho11, Goe79, Gol83a, Gol85a, GPR81, GG11b, GB77, Gra89, GJ84, GM14, Har14b, Har14c, Hec19, Hec16, Hog97, Hol78, Hol83b, Hol82, Hol83c, Hop77, Hop84, Int78, Isa89f, Isa90a, Isa91d, Isa91h, Isa91r, Isa96d, Joh86, Kah86, Kap82, Kaz14, Ken87c, Kla88, Kok88b, KK19e, KN12b, KN13b]. weapons [KN14b, KN14f, KN17a, KN17e, KK21a, KK21b, KK21c, KK21d, KK21e, KK21f, KK22b, KK22a, KK22c, KK22d, KK22e, KK22f, KKR23a, KKJK23a, KKJ23a, KKJK23b, KKJ23b, KKR23b, KKJK24a, KKJK24b, KKJK24c, KKJK24d, KKJK24e, KKJK25, Lan14, Lan16, Lat91, LE85, Len20, Loc91a, Lun79, Mal14, Mar83b, MK02a, May85, Mec20a, Mec20h, Mei20, Mey83, Mey85, Mia16, Mil10, Mil12a, iM16, Moo96n, Mor83c, Mor84c, Mor85b, Nel81, Nie14, NH08b, NA87a, NA87c, NA87k, NA88j, NA88t, NA88q, NA89q, NA89p, NA89r, NA90i, NA90h, NA90j, NA90k, NA90n, NA90r, NA90u, NA91p, NA91o, NA91q, NA92p, NA93s, NA94c, NA94g, NBF94, NA95e, NK10b, NK11, NK16, OGP84, Oel14, Oel20, OW90a, PC92b, Pan92, PE24, Per91b, Per20, Pin24]. weapons [Pod16, PN09, Raj14, RK19, Rap11, Ray90, Rob78c, Roc76, Rof16, Rom83, Rom92, Ros87a, Ros94, RE19, Rot77, Rot97h, Sai14, San16, Sar83, Sau10, Sau17, Sch78b, Sch98, Sha87b, Sha18, Sha82, She16, Sho80, Sho92a, Sin14, SOE21, Smi90b, Smi92a, Smi93d, Sol90, Spi83, Bul09h, Ste86a, ST22, TBBS85, Tay86, Tra24, Tsi77b, Tsi80, Tsi82d, Tsi83d, Tsi83e, Tsi83i, Uni78, Ust93b, WH89, WSG^{+}15, WP19, WF86, Wei84, Wel16, Wes85, Wil77c, Wri87, Wri89, Yor82, Yor88, ZRG83, ZC92, Zik21, Zwe84, Zys17, Ano78n, Ano86i, IK21, Mil85b, NA93o, Ano78f, Bro61, For83, Fre84b, Ham93, Mol62a, Mos81a, Sei71b, Sid69, Han94, Gom84a, Ste80a]. weapons-grade [Rot77]. wearing [Kim95, NA89c]. Weart [Smi88, Bul08o]. Weather [Jac73g, Par73, Sch70a, TT69, Bal79c]. weave [Rot94f]. Weaver [Gru70b]. Web [Bro67b, Moo97c, Moo97f, Rot94f, Lor99, NA01]. Webber [Gri83a]. Wedge [O’L75]. wee [Moo96e]. Week [Ber76b, Kal82a, Rab51j, Ano48g]. weekend [CR81]. Weeks [Ano51t, Ano53x, Ano53-47]. Weighing [Bin04, Mil12b]. Weight [MV04]. Weiler [Div79]. Weimar [Kin69]. Weinberg [Cha79a, Moo96c, Mot68, Tsi72a, WI71]. Weinberger [Ano98l]. Weiner [Pal64b]. Weintraub [Ano85d]. Weir [Pla57b]. Weird [Blu88, Rot02c]. Weisberger [Bus85]. Weiss [Fel81m, Gab73, Bal79c]. Weisskopf [All56l, Bur48, LL90]. Weissman [And83]. Weiszäcker [Bar79b]. Weizmann [Cor53]. Weizsäcker [Bur65, Haw50]. Welch [Leb94]. Welcome [Ano53-48, Ano63w, Ano94l, Arn92d, Har93, Ano98o, Cop80, Rot96f]. welcomes [Ano97p, Ano98j]. Welfare [Ano46-53, Eri95, Lon50, Pla73, Sab72]. We’ll [Pan00, Ano93r, Ano97l, Dow84c, Moo94g, She85f, Wei83c, YES^{+}21]. Wellerstein [Eav17c]. Wellhead [GW02]. Wellsian [Wel51]. Wellstone [Mor82a]. Wendt [Sne57]. Wenk [Woo73b]. Went [Ber76b, Bou74, Kla19, Lor00a, Osn77, Wil79]. We’re [Ano93r, BR57, JG96, Rab59p, AH91c, Dys75, Moo01f, NAB99b, NAB00, Ano90-33, Jr22c, Thu79, Alb98b, Ano90-32, Ano92r, BP95, McK25b, NA90p, Zan91]. weren’t [Coo76a, Ide93, McM94b]. Werner [Hei54, Swe93, Cas92]. Wernher [Neu07]. Wertheim [Ano80l]. Wes [Wun82]. West [Ano89g, Ard93b, Bir87, Bou99, Bro57b, Cav57, Ein48a, Har92b, RK89a, Sol80a, The89a, Mot89, Ano79c, Ano84e, Ant78, Ber65, Blu84b, Brz82, Cha89b, Jr22a, Fre77b, Foz61b, Hil16, Hil20, Hol85a, Ken81, McC88b, Nee79, RK87b, Roc78, RHKK78, RF65, Sak78, SW59, Sha93c, Som84d, Sut01, Thi58, You64a, vL65, vL85, vMK88, Bar88, Nic54a, Ten75, Wri63a]. Western [Ano57d, Cra84b, Els90, Gué70, Kap95, Mul66, Ray81, Sak69, Som91, Stu89, Swe86, Wri63a, Aga79, Ano49e, Ano60j, Ano85y, Ano98m, Aro60, Buc60, CB95, Daa87, Has88, Hay55, Jac69, Kis94, Kno60, Koe77, Mül90a, NA90h, Oli48, You63b, Sch69a]. Westgren [Kat52b]. Westing [Kri85, Mye87a, Ano91q, Kri85]. Westinghouse [Ano49-41]. Weston [Som84c]. Westward [Str93]. Wet [SSC68, Spi64]. We’ve [Dow80e]. Wexler [Dow83d, Dow79e]. WFUNA [Ano49-42]. whacked [Isa92d]. Whale [McV71, O’C87]. whales [Ano94m]. Whatever [Cic74, Dee59, Gay98]. Wheat [Her67, Rab67c]. Wheel [DeB81a, Jan76a]. wheeler [Ano90-46, Ano68g, Phe63]. wheeler-dealers [Ano90-46]. Wheeling [OW99]. Wheels [Tor77b]. Whelan [Zel92]. Where [Ano63l, Ano90-33, Ano91-38, Ano96j, BW18, BSS06, Cra84g, Cum97b, Dar91, Das66, Dro23m, Eve89, Fel78j, Gom84b, Gor96c, Hor71, Isa89f, Isa00f, KK12, Kre92b, Kre89c, Kri75, Lew69b, Lil49d, Lon67, Lon73, Moo01f, Mot51, NA89c, NA91q, NA92m, NA97f, NAB99b, NAB00, NK06e, Pag61, Rab51i, Ran82, Rob91b, Rot00c, Spa69, Spi74, Sto65b, Wei17, Wil69a, vH01, Cam95a, Dro21d, Dro23f, Mac11, Moo95o, Rap65a, Lor00b, Gle52]. whereby [DS47b]. Wherefore [Don86b]. Where’s [Moo94i, Wei94b]. Wherry [Ano48n]. Whether [Kri75]. Whew [Ano90-34]. Which [All59l, Ano91-39, Hib95f, Kou75, Rab46f, Sch47, Hec22]. While [Ber01, Lei74, Ros06a, Wig61, Ano94j, Ano97k, Mec20g]. Whispers [Tum87]. Whistle [Ano94j, Tuc77]. Whistle-blowing [Tuc77]. whistleblower [McG90, vH93]. Whistleblowing [Fel83d, Klo85]. Whistles [WF77]. Whitacher [Dow80g]. White [Hor77, Moo94a, O’L73, Pas76, San57, Zei54, Abb91, Bin79, Dro19, Ano87p, Cla66, Dow80i, Fla78, Jac74a, Lal66e, Moo98h, Wol91b]. Whitehead [SH89a]. Whiteside [Tho80c]. Whitfield [Boy91]. Whither [Ano97m, Bar80a, Mei77a]. Whiting [Kli90]. Whitman [Mec19a]. Whitney [Kat53a]. Whitrow [Smi58f]. Whittaker [Lit54]. Who [Abb84b, Ack92, Ano61w, Ano61x, Ano84a, Ano88w, Ano91-33, Ano91-42, Ano93u, Ano98p, Ano99-30, Bar83b, BM93, Bru59, Cha85b, Cha68b, CC13, Cle01, CE88, Dow84a, Eav16, Gar83b, HFMH89, HM90a, HM90b, Isa88a, Isa89g, Kal95a, Lea07, Mar69, NA92j, Pen84, Rab50c, Rat62, Rei91, Ros68b, SG83, Sch80b, Sch88b, Thi53, Tho88, Tra90, Van60, Wei97, Ack86, Ano90-52, Ano90-61, Ano96j, Ark89b, Gau24, Man80, MRH16, Per88b, Rot93e, Sha92b, Spe87a, Sto19c, Wri77, Aug55, BES02, Lei82b, Ano91-33, Dow81i]. Whole [Gle94, Ric06, Ano15e, Bas80]. Whole-earth [Gle94]. Whom [Ben72, Ois72, SvH77, Ack86, Ano88w, Gor95d, Ano53-48]. Whoops [Rot94j]. Whose [Ano86h, Ano90-51, Ark02d, Fra58, Gre05, Rot98c, Dow83e]. Whyte [Cro67, Dro24a, Lan76, Smi59b]. Wichita [Gol84b]. wide [Bre13, Gre84a].

Widening [Ano95i, Bla69a]. Wiecek [Abb85]. Wiener [Mei68]. Wieseltier [Sha84b]. Wiesner [Cle95]. Wife [Wei94a]. Wigley [Ano14g]. Wigner [Fla68, BJH68, Kue61, Wig66, WF68]. Wiktor [Ran85]. Wilcox [Dow86, Pol91b]. Wild [Lew73a, Bou99]. Wildavsky [Hol83a]. wildfire [LBCA23]. Wildlife [CKA10, MM11]. Wilkinson [Gri83a, Nei82a]. Will [Ano50s, Ano95n, Ano97m, BBZ01, Cah75, Cha68b, Cle89, Com74, Dro23j, Gei17, Gol86a, Ham91b, Hit03, Isa85f, Kri75, Lem81, Lew69b, McV71, Mue01, Mul52, Mul68, Nan71, Pra03c, Rab59a, Rab69d, RK89b, Sal16, Say84, Sok20, Spi59, STRS80, Sto23, Tha00, Ure48c, Vis56, Wan05, Wee98, Wol91b, vB67, vH79, Ack92, Ano15e, AH22, Bar77b, Dro22e, Eav16, Gho15a, HBD17, Hec22, Kap91b, Kel90, Khr97, KGS25, Kre92b, Men91c, Moo92a, NH08b, NA94e, Nov77b, Nuc18, Pol91a, Rot77, Rot93e, Sau17, Sch18, Sez20, Ste77c, Zal19, Mec22c, Sto58]. Willard [Mat56, Pre62]. Willens [Dow79e]. Willful [Ves05]. William [A.52, Abb85, Ano78d, Ano78k, Ano86k, Ano87q, Bal79c, Bla84, Cal84a, Cra83b, Div79, Don89, Ehr81, Fel83e, Fra83, Gai92, Gol88b, Gol67a, Gre55a, Han92b, Han98a, Haw58b, Her88b, Ing82, Ing56c, Jac80b, dK60, Lac85, Lap57, Leb91b, Lev84b, Low93, Lyt79, McC96, Mec19c, Met83, Mil86b, Moo95c, Mor87c, Pap88, Pla57b, Rep86, Rob69, Roc85, Rod70, Rot93a, Sch68e, Sco74, Sim85, Ste76a, Sto55a, Sto58, War49, Wou81, Ano88x, Ano88y, Ano22c, Ein53, Kee56, Ure56, dF79h]. Williams [Hol62, Hur78a, Kre89c, Qua80, Sti83, Wil82b]. Williamstown [Lew70e]. Willis [Dow79c]. Willrich [Ano78n, DeN74a, Fre76b]. Willy [Ano22c, Bra87]. Wilson [Ber73c, Hol83a, Tho80g, Ano65f, Ano85h, Dow80d, Sch78a, Sil71]. Wilting [Sha99]. Win [Fis62a, Fra80, Mor87c, Ray89, Rus58b, Ack88e, Eav16, Isa88e, Isa92d, Isa97b, McK23a, Rus20]. Wind [Gre68b, Ing75, Ing79b, Mes83a, Par73, Sør76, Wee03, Wea99, Ano93w, Com76a, DJ13, Dro21c, Zel76a, Mer79]. window [Car84b]. Windowshopping [Ded59]. Windpower [Tor77c]. Winds [Ack87i, Dow81c, Ano93z]. Windscale [Ano94o, MP78, Pat78c, Woh78]. wing [Isa95a, Isa95e, Isa96d]. Wingo [Fri64]. wings [Ano97k, Rot93r]. Winkler [Abb93]. Winner [Ark00e, Ano90-30, Ano96j, Ano05a, Dro23a, Aue78, O’C87]. Winners [Ano82r, CCD^{+}52, Dow83d, Sch69c, Dow79e, Dow82h]. Winning [BG14, Esp14, Eva04, Her82a, Hun14, Tom14, Dro17, Fie24]. Wins [Isa03e, Ano86n, Ano87u, Ano87z, Ano91u, Ano92l, Ano93-28, Isa86d, Sch88b, HD82]. Winslow [Rab98]. Winston [Chu50]. Winter [Ehr84, Rob91b, HH87, Kap85, Mal85, Nat85c, Pos85, Rob89, Tur17]. winters [Dro20b]. WIPP [Ial18, Kla19]. wire [Ano95m]. wires [Ano78p]. Wis [Jac71b]. Wisconsin [Day95]. Wisdom [Aue05b, Gol07, Mon84, Kel87, Lif76]. Wise [Thi52]. Wiseman [Dow80h]. Wish [Sti57, Rot04h]. Wishful [Cav07]. Wishing [Ste76a]. Wissenschafter [Ano68l]. Wit [She74b]. Witch [Ano99-31, Ure49a]. Withdraw [Cra67b]. Withdrawal [Men92b]. Within [Ano47n, Coo06, Ker06, KAT^{+}71, Tsi75a, Wyl69, Coh86, E.77, Har79a, Wal21b, dF79a, Lal70]. Without [Ack09, Ano78o, Bor52a, Dow79e, Dow82h, Dow83d, Fed59a, FW64, Han95b, Kon67, Mar64a, Mar48b, Rab73e, Rob75b, Sto59b, Was64, WB51, Zha99a, Alb87b, Ano51x, Ano52u, Ano94k, Bla88, Cha79b, Cla59, Fin90, Har88d, Isa96b, MCC51, Mor59a, O’B82, Pat78a, Pos85, Rab56b, Roc81, Som84b, Ure48a, Vis59, WP19, WF90, Wil77c, vD78, Rab73d]. Witness [Les94]. Witnessed [DeB81a]. Wittner [Mor85a, Olu71, Win94, Win98]. Wizards [Rod85]. WMD [Bou11, Car13, LST10, Mey18, Tay21, WL21]. WMD-free [Bou11]. WMD-related [LST10]. woes [KSS17, NA90q]. Wofford [Isa92d]. Wohlstetter [Woh78]. Wolf [Spi54, Sti83]. Wolfe [Roc85, Wol50]. Wolff [Dev69]. Wolfgang [Kon67]. Wolfle [Ano60f]. Wolfowitz [Uhl03a]. Wolfsthal [Mec19f]. Wolton [Hal93a]. Woman [Sch72d, Bet72]. Womanpower [Rab51d]. Women [Ano46-50, Ber66b, Cha82c, Dje74, DB67, Eat62, HC83, Kaz14, Mal14, Nie14, PB68, Sin67a, Sin14, Ano93-32, Azh95a, Gya14, Har14a, Joh94d, Mar76, Rot91a, Sha14, Ano89f]. Won [Ken60b, Moo91g]. wonder [NA89f]. wondered [Ano94k]. wonderful [Moo95r]. Wonderland [AK93]. Wondrous [Spi64]. Won’t [You75, Bro85e, HS13, Kre85b, Mei20, Men99, WLP97]. Wood [Dro22i, You77b, AGB22, SMRVS22, Smi60]. Wood-burning [Dro22i]. Woodhouse [Wri63a, Per87b]. Woodley [Kna54]. Woodruff [BJCP88]. Woodward [Har88c, Sha91a]. Wooing [Har90e, Mir89, Tod81]. Woolf [Sto19a]. Woollacott [Tic94]. Woolsey [Pra93b]. Woops} [Ano93t]. Worcester [Wyl56]. Word [Ano92q, Fly06a, Gre73c, Ano91-36, Ano92t, Ano95i, Ark94a, Ark94d, Ark94b, Ark94c, Ark95a, Ark95c, Ark95d, Ark95b, Ark95e, Ark96d, Ark96a, Ark96b, Ark96c, Ark96e, Ark97c, Ark97b, Ark97f, Ark97d, Ark97e, Ark97g, Ark98a, Ark98c, Ark98b, Ark98d, Ark98e, Ark98f, Sco08b]. Words [Fel81l, Sta60a, VT73, Ano96o, Ano98n, Ano12a, CKN^{+}94, Moo96m, Moo98d, Rot91a, Dow84d]. Work [Ano49-35, Day75f, Day88, For02, GF73, Hol97, JCKG^{+}50, Koc98, Lau75, Mal50, Mel58, Men99, Pop52, Rab59a, Ros06a, Ros08, Ski48, Ure49a, Wel77, ZvH01, Alv01b, Ano88g, Ano91-42, Azh95a, BWL21, Bro85e, Dro16a, Far77, Fra76a, Jan81b, JS97, Lev89, LP10, LC93, MJ80, Rot93f, Rot94c, Sam18, Smi93b, WLP97, Woo98a, YES^{+}21, ZG83, vL47]. workable [Pap98]. Workbench [Bul07a]. Worked [Kre89a, Ano90-52, Mül89, dF79d]. Workers [Ano46-108, Kan09, Obe74, Con90b, Rot78b, Wal21b, Ano46-30, Wol55, Sim84b]. workfare [Isa92g]. Working [Ano49q, Ano60a, Ano60c, Ano60d, Hal84, Ing52a, Rab47o, Ste81b, Ste82b, Veg99, AO95b, Dro22h, Joh94d, Mey18, Mot51]. workplace [Ano84j, Sim84b]. Works [Ber65, Gab73, Ano88l, Cul14, Gla85, Kap81b, Pai80b, Ano55i]. Workshop [Sch78a, Ano85x]. Workshops [Lap57]. World [Amr54, Ano46u, Ano46-108, Ano47-29, Ano52-27, Ano54t, Ano58c, Ano64p, Ano64q, Ano70a, Ano74i, Ano78n, Ano78m, Ano87-29, Ano90m, Ano93l, Bar76d, Bar78b, Bar79a, Bar80b, Bar81a, Bar74, Ber72d, Bla82b, Bla57, Boh50, Bos05, Bro79b, Bro47e, Bro47f, Bro56c, Bro58, Bus85, Cah76, Car49a, Cav47, Cha79b, CLR06, Con64, Con67, Coo56, Coo62a, Dea51, Den47, Dre51, Dub59, EE86, Eva63, Fel75b, Fel75c, Fer57b, FP05, Fis83, Fit51, Fol74, Fox46, Fre76b, Gen84c, Gid73, Gol88a, Gom75a, Gom75b, Gom75c, GR63, Hal75b, Har01, Hoe88, Hor89, Hun87, Hur78a, Hur78b, Int67, Ikl75, Ima71, Jac65, Jac66, Jac73a, Joh89, Joh97a, Joh00b, Kam71, Kin67, Kre89c, Kre98, Kus68, LaF90a, Lea47, Leg57, Lei72b, Ler80a, LT70, Lev72a, Lew56, Lov74b, Lov74c, Mar64a, Mar78a, Mar98c]. World [May47b, McG84, McM48b, Mer90, Mey49a, Mic68, Mil68, Moo95d, MW47, Nad98a, Nas04b, Nie49, Nie63, Nin59, NA99e, Not51, Opp48b, Orr56a, Orr56b, Pau57, Pie68, Pow90a, Pow91b, Pri74, Rab47f, Rab51m, Rab56k, Rab56i, Rab58a, Rab63g, Rab64e, Rab69f, Rab70c, Rao77, Rea80, RMW54, RP66, Ric06, Rig65, Rob75b, Roc85, Rus56a, Rus58b, Sak75, Sal64b, Sal68, Sal70, San68, Sch67b, Sch08b, STFK09, Sch68h, Sea73, Sem64, Sha54b, Sha59, Sim54d, Siv75, Smi85d, Sot67, Spu73, Bul02e, Bul09a, Ste95, Sto58, Str54b, Str86, Sul58a, SG69b, Szi47a, Tay47, Tel48a, Tel76, Tru85, Ure46b, Ure46a, Ure48a, Usb47, UH56, Val86, VFIS63, Wad63, Wag61, War68, WM82, Wei56a, Wei68c, Whi79, Wil75b, Wil71, Wil84, Win94, Win98, Wis55, Wri48, Wri63a]. World [Yut69b, dF79b, vL48b, Abb12, Ack91d, Aja78, Alb98c, Ald17, Ano76b, Ano91-67, Ark94d, Bar78a, BWA85, Bro79e, Cas19, Cha75d, Day77o, Dea91, DJ13, Dw77, Dow91, Jr22c, Dro22c, E^{+}84, Fah91, Fel76e, Fel78c, Fel81e, FS95, Fie19, For88, Gar89, Gor96c, Hol96a, Hui91, Hun85c, Isa88c, Isa90o, Isa93p, Ive16, Joh91, Kap91b, Kre91a, Len20, Lym11, Mac77a, Man77, Mat80a, McK23b, Men91a, Men93b, Men91d, Nas18, Nea78a, Nea78b, Pug77, Par17, Per91b, Pla78, Pol91a, Rep10, RK91, RF21, Rot77, Rot96f, Rus20, Sag83, Sca14, SFT11, SF12, SF14, Sch77b, She94, Sok91, Sol91, Ste91c, Ste81a, Tho81b, Wil77a, Wio77, Woo91, Yor75b, You91, Zhe90a, vD78, vMC98, Ack88l, Ano52g, Ano86i, Ano89c, Ano90-64, Ano96-31, Ark85f, Ark98d, Com48a, CF74]. World [Con90a, Coo71, Cra84f, Dha90, Ein48a, Ein79a, FKM87, Gol69b, Goo86, Got77, Gou75b, Har87c, Hay78, Kal83a, Kar88c, Kar88d, Kla84, Kla90, Lev70, LGW92, Mak76, Sen47, Moo96n, Mul53, Pow89a, Pri89, Rab51j, Rav84, Rea80, Red78a, Red78b, Rev92, Rya89, SR80, Sal88, Sha59, Sto59b, Ven79, VSD^{+}09, Win90, Ano78m, Ano91q, Bur65, Eas85, Fin65, Hug57, Ker68, Moo93b, Moo98b, Pat82a, Ray81, Sno68a, Wou81, All55b, All55a, Ano46v, Ano86l, Ano87r, Ano88d, Ano90p, Ehr82b, G.84b, Goo86, Mei52, Mye88, Rab73e, Rot93a, Sto59b]. Worlds [Ber75b, Cha63b, Key66, Kre86c, Tho80d]. Worldwatch [Tod78a, Tod79]. Worldwide [E^{+}84, Fel83a, KN14f, KN17e, NA90e, NA93h, NA96c, NK09b, Hes78, Mec19e, NA89h, NA91f, NA92g, NA94b, NA95c, NA98c]. Wormall [Wol55]. worries [McC86a]. Worrisome [Raj14]. Worry [Hir02, Ano91-48, Ano95j, Ano98p, CW23, Kle81]. Worrying [Boy05, Ber23, ZA04]. Worse [Nad04, OM97, Pai02, Woo97b]. worsening [Mec22a]. Worst [Hur77b, Kri75, Rab66c, Alb94b, Bad78, Kla82, Leg92b, Pat86c]. worst-case [Alb94b]. Worth [Che52, Gol74, Lev77c, Mad63, Mel75, Rat69, Sim60a, Ano90-49]. Worthless [Rob78a]. Worthwhile [Spi57b]. Would [AS02, Ano65h, Che52, Fet87, Gei16b, Gla75a, MHD^{+}05, Rub74, Rus53, STC74, AB18, Cha84, Col79c, GLH^{+}79, Gol88f, Hay79, Isa88a, Isa95e, MSW20, Mil79, PD89, Rot92g, Sha84d, Tan88]. woulda [Ano98p]. Wounded [Lar99]. Woydyło [Ran85]. wrap [Eps95]. wrap-up [Eps95]. wraps [Ano91-59]. wrapup [NA88a]. Wrath [Bec78, Rom83, Tsi83g]. Wray [McK23b]. Wright [Ano91s, Ano46-80, Ano87r, Bro57e, Mes64b, Wri64]. Write [Eps92b, Woo85, Haz82]. Writer [Asi62b]. Writers [Ano63-29, FU63, LaB81]. Writes [Con48, Ano98n, Man80]. Writing [Ano63-29, FU63, Rot01b, Zel75, Dro23j, Yev86]. Written [Ano63-29, Dow80f, FU63, Ano91-63, Kat52b]. Wrong [Ant74, Bou74, Gar00b, GH94, Kir91, Sau01, SH89a, Ump75, AA54, Coh95a, Cra98, Cum15, HI94, Isa95e, Kla19, Lev87, Osn77, RGK95, Rot97f, SB89, Tsi82b, Wil79]. Wrongheaded [Aft99]. wrote [dW90a]. wrought [Law75]. Wulf [Han76]. www.thebulletin.org [Ano08a, Ano08b]. Wylie [WKW57, Wyl57]. Wynen [Sei71b, Sei73]. Wynner [All55b, Ano46-80].

X [Ano54f, LD79, Toy96, Ano86-32, Blu88, DeW88, Gho15b, Ken87b, TS92, Uni57]. X-37B [Gho15b]. X-out [TS92]. X-ray [Ano86-32, Blu88, DeW88]. X-Rays [Uni57]. Xeroxing [Fre72b]. XI [Ano55d]. XII [Pop52, Ano56e, Ano64p]. XIII [Ano57e, Ano64q]. XIV [Ano58b, Rab65d]. Xue [May89]. XV [Ano59b, Ano66p]. XVI [Ano60g, Ano67l, Rab67h]. XVII [Ano61e, Int67]. XVIII [Ano62m]. XX [Ano64h]. XXI [Ano65d]. XXII [Ano66a]. XXIII [Ano67c]. XXIV [Ano68b]. XXIX [Ano73a]. XXV [Ano69b]. XXVII [Ano71a]. XXVIII [Ano72d]. XXX [Ano74a]. XXXI [Ano75b]. XXXII [Ano76a]. XXXIII [Ano77k]. XXXIV [Ano78h]. XXXV [Ano79b]. XXXVI [Ano80e]. XXXVII [Ano81d]. XXXVIII [Ano82b].

Y2K [Kra99]. Ya [Ses83]. Yager [Jac80b]. Yakovlev [Ekl80]. Y’all [Rot93u]. Yalta [LaF62, Fel77o]. Yamamoto [Che88]. Yamazaki [McC96]. Yancey [Sch68e]. yanked [Woo97d]. Yankee [Ano87n, Woo97d, Woo01a]. yanks [Day77b]. Yardstick [Nie66a]. Yarmolinsky [Phe56, Sco71]. Ye [Wal74]. yea [Ano89i]. Year [Ano55m, Asi62a, Ber73a, Bro62b, Isa91a, Jac73f, Lew65b, Mei68, Nor63, NA88e, Rab48e, Rab57e, Rei63, Sim47, Str05a, Tel48b, Win84a, Ano87-30, Ano88v, Ano92t, Ano92q, Ano14c, Ash60, Bar85c, Bro86c, Dje81, Eps12, Fel76f, Fel82l, Hib12, Isa84a, Isa90a, Isa98e, MD24, Pai82i, Pos21, Spe81, Spr19, Wei94c, Fed59c, Rab53b, Rab58c, Rab62g, Rab63c, Rab64d, Rab65f, Rab66d, Rab67g, Rab68b, Rab69e, Rus55c, Sim57a, Sti57, Tro58]. Yearbook [Ano86l, G.84b, Ano78n]. Years [All53, And65, Ano48q, Ano48c, Ano63m, Ano66f, Ano90i, Ano91j, Ano91k, Ano96a, Bal54, Bec75, Can63, Deu67, Fra54, Gre67b, Gui65, H.85, Her83a, Lal67e, Lap53, Lew56, Mak03, Moo95e, Mul65, Nas03b, Nei82a, Nel75, Pie54, Rab47m, Rab51b, Rab52c, Rab56i, Rab70e, SA62, Bul02a, Str06d, Way51, Wie48, AO95a, Ano77a, Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77g, Ano77h, Ano77i, Ano78a, Ano78b, Ano89q, Ano89z, Ano89y, Ano90-56, Ano90-57, Ano90-62, Ano90e, Ano90f, Ano90h, Ano90c, Ano90g, Ano90b, Ano90d, Ano91e, Ano91f, Ano91d, Ano91b, Ano91g, Ano91h, Ano91c, Ano91a, Ano91i, Ano91l, Ano92i, Ano92g, Ano92j, Ano92c, Ano92d, Ano92a, Ano92b, Ano92e, Ano92f, Ano92h, Ano93a, Ano93e, Ano93f, Ano93d, Ano93h]. years [Ano93b, Ano93i, Ano93g, Ano93c, Ano93j, Ano94g, Ano94c, Ano94d, Ano94a, Ano94e, Ano94b, Ano94f, Ano95e, Ano95f, Ano95b, Ano95g, Ano95d, Ano95c, Ano96-28, Ano96c, Ano96e, Ano96d, Ano96b, Ano97e, Ano97b, Ano97f, Ano97a, Ano97c, Ano97d, Ano98b, Ano98e, Ano98c, Ano98a, Ano98f, Ano98d, Ber84b, Cra87, Dow82i, Jr16, Fel85a, Gom82b, Gor11, Gor20, Isa95e, Jon90a, Jos87, KW15, Kap82, KSS17, Man77, Mar98a, Moo97k, Nor94b, Pat81c, Rab20, RH14, Rot81b, SW80b, Soc20, TvH17, Tsi82d, Was82, WB95, Wit95a, Yor75b, Yor82, Zag95e, Bas80, Gre54, Leb91a, Pla57b, Rau91, Rob69]. yeas [Isa88c]. Yehuda [dF79f]. Yellow [Col84c, Kal82i]. Yeltsin [Arb93, CNB95]. Yergin [Fel77o]. Yes [Ano93u, Cri94, Eve88b, Joh01, RMW54, Rot95l, Rot96c, Ano11g, BT95, Isa95d, Mar98b, Ehr75, Rab73g, Ano95r, Fei71, Kah86]. yesterday [Thu79]. Yevgeni [Bou87]. Yevgeny [Yev86]. Yevtushenko [Yev86]. Yi [Fra81]. Yi-fu [Fra81]. Yield [Ibs87, Szi49a, Szi63a, Zik21]. Yields [Bar64a]. Yih [Pre96]. Yo [Mar65i]. Yo-Yo [Mar65i]. Yonah [Ano87o, Nei82a]. Yongbyon [Alv03, Alv20]. Yoo [Mye88]. York [Ano84k, Lef77b, Lev51b, Rot97c, Spi50, Tel49, Wil71, Ano60f, Ano81a, Kis81, Mar65e, Mar66b, Nel74b, Zac76]. Yorka [Ano92x]. Yorker [DeB81a]. Yorkers [SB23]. Yossi [Mos91]. Yost [Sch81a, Stu89]. Youl [Mye88]. Young [Bal79c, Key81, Lan90n, SS69, Bar81i, Fri20, Kli80, Lem81, Mec15a, Med81, Wil85b, Ano84o, Fel83e, Jas81, La 91, Lal67a]. Youngest [Ano89m, Ano98s]. You’re [Ama03, Fel83d]. Yourself [Cal61a, Lor01a, Bul01a, Cho81, Fly99a, Ros90b]. Youth [Ols16, Rab69k, Sie72, Wil75c]. Yuan [Ano87-32]. Yucca [CP98, CBSF95]. Yudkin [Gol93a]. Yugoslavia [Ano63o, Sui71, Sup61, Woo93]. Yukawa [Ten75, Toy82]. Yuki [Joh97a]. Yuri [Ano78q]. yurt [Moo93d]. YV [Ano78n].

Z [Ano78n, Moo94a, Ber23]. Zacharias [Fel92]. Zachary [Ano98g]. Zacher [Jas81]. Zaelke [Dro19]. Zagalsky [Zag94f]. Zaiisk [Dow84d]. Zaire [Ano12d]. zap [PGL91]. zapped [Isa88c]. Zapping [And78a, Lou90]. Zarf [Bla02]. Zaslavskaya [Kap91a]. Zbigniew [Her83a, Mul65]. Zdorovye [Isa92i]. Zealand [Amb01, Cle87a, Cle87b, Cle92, Sal87]. Zehner [Ano12e]. Zeit [For80]. Zeman [Boy05]. zen [Fra91]. Zero [Ano82o, Ano89g, Bea89, But00, Chu94b, Dow82d, Dow84b, Kal82a, KS94, Ken07, Ack87e, Al 15, Ano95n, Ano97k, Bin15, BV10, KR11, Moo93b, Rep10, Suz11, Ulg15, Zam98, Dow81d, Dow82h, Kal83c, Ano98h, Pol93]. zero-risk [Suz11]. zero-sum [Ano97k]. Zero/Phase [Ano82o]. zeroes [Ano98k]. Zerwick [Wil68b]. Zhores [Ano78e, Hec70, Par80, Rab72a]. Ziman [Sno69]. Zimock [Kre88a]. Zinberg [H.84a]. Zinn [Fra91]. Zion [Gus70]. Zirconium [Pal76, O’N95]. Zonal [Ore63, Soh62]. Zone [Bar66b, BS01, Com75b, Sie05e, Sti63, And85, Fry85, Har14b, Lan14, ML82, RP82, Sai14, Zei85, Woo97e]. Zones [Das68, How62, Don85]. zookeeper [Her90a]. Zoonotic [Kah07]. Zorin [PT87, Ras83]. ZPG [Mul74b]. Zraket [For87]. Zuckerman [Lak66, dF79f, Her90a, Nad79, Tuc72, Kil78]. Zuckert [Bro57a]. Zühlsdorff [Wri63a]. Zukunft [RK87a]. zum [For80]. zur [Ano68l]. Zweig [Con67].

References

ANO:1983:A

[A N83] A Nearby Observer. Afghanistan. Bulletin of the Atomic Scientists, 39(6):16–23, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author is Charles F. Dunbar: see addendum [Ano83a] and comment [Ege83].

A:1951:TPU

[A.51] M. A. Technical programs in underdeveloped areas. Bulletin of the Atomic Scientists, 7(7–8):223–226, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

A:1952:BRB

[A.52] M. A. Book review: Mr. President, by William Hillman, with photographs by Alfred Wagg. Bulletin of the Atomic Scientists, 8(4):125–126, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

A:1954:BBB

[A.54] F. L. A. Books: Radiation Biology, Volume I: High Energy Radiation, edited by Alexander Hollaender. Bulletin of the Atomic Scientists, 10(10):394, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

A:1955:BPN

[A.55] H. A. The British program of nuclear power. Bulletin of the Atomic Scientists, 11(6):224–225, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

A:1962:TPC

[A.62] R. A. Tenth Pugwash Conference. Bulletin of the Atomic Scientists, 18(9):39–40, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alsop:1954:RW

[AA54] Joseph Alsop and Stuart Alsop. [On the right to be wrong]. Bulletin of the Atomic Scientists, 10(5):178, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1952:PUC

[AAF^{+}52] Herbert L. Anderson, Samuel K. Allison, Enrico Fermi, Willard F. Libby, Joseph E. Mayer, Edward Teller, and Harold C. Urey. Protest from University of Chicago scientists. Bulletin of the Atomic Scientists, 8(7):255–256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Abrams:1977:MBA

[AB77] Nancy Ellen Abrams and R. Stephen Berry. Mediation: a better alternative to science courts. Bulletin of the Atomic Scientists, 33 (4):50–53, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pas77].

Albright:1989:TF

[AB89] David Albright and James Beard. The tritium follies. Bulletin of the Atomic Scientists, 45(9):42–45, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mar90].

Arnold:2005:TNE

[AB05] Lorna Arnold and Andrew Brown. Time for a nuclear entente cordiale. Bulletin of the Atomic Scientists, 61(5):76, September/ October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
5/76.

Acton:2008:NNN

[AB08] James M. Acton and Wyn Q. Bowen. Nurturing nuclear neophytes. Bulletin of the Atomic Scientists, 64(4):??, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/4/27.

Arguello:2018:GIT

[AB18] Irma Arguello and Emiliano J. Buis. The global impacts of a terrorist nuclear attack: What would happen? What should we do? Bulletin of the Atomic Scientists, 74(2):114–119, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1436812.

Abas:1989:RFB

[Aba89] Bryan Abas. Rocky Flats: a big mistake from day one. Bulletin of the Atomic Scientists, 45(10):18–24, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1946:RSD

[ABB^{+}46] L. W. Alvarez, R. F. Bacher, M. Benedict, H. A. Bethe, A. H. Compton, Farrington Daniels, J. R. Oppenheimer, J. R. Ruhoff, G. T. Seaborg, S. H. Spedding, C. A. Thomas, and W. H. Zinn. Repercussions of the State Dep’t report: Denaturing fissionable materials. Bulletin of the Atomic Scientists, 1(9):11, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1950:LUP

[ABB^{+}50] Samuel K. Allison, Kenneth T. Bainbridge, Hans Bethe, Robert B. Brode, Charles C. Lauritsen, F. Wheeler Loomis, George B. Pegram, Bruno Rossi, Frederick Seitz, Merle A. Tuve, Victor F. Weisskopf, and Milton G. White. Let us pledge not to use H-bomb first! Bulletin of the Atomic Scientists, 6(3):75, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1979:RWT

[Abb79] John Abbotts. Radioactive waste: a technical solution? Bulletin of the Atomic Scientists, 35(8):12–18, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1982:ISS

[Abb82] John Abbotts. An industry in search of subsidy. Bulletin of the Atomic Scientists, 38(7):47–51, August/September 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [MA83].

Abbotts:1984:BRB

[Abb84a] John Abbotts. Book review: Beyond Dumping: New Strategies for Controlling Toxic Contaminations, edited by Bruce Piasecki. Bulletin of the Atomic Scientists, 40(10):28–30, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1984:WPR

[Abb84b] John Abbotts. Who pays for radioactive rubbish? Bulletin of the Atomic Scientists, 40(7):24–38, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pen84].

Abbotts:1985:BRB

[Abb85] John Abbotts. Book review: Nuclear America: Military and Civilian Nuclear Power in the United States, 1940–1980, by Gerald H. Clarfield and William M. Wiecek. Bulletin of the Atomic Scientists, 41(5):43–44, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1989:NPA

[Abb89] John Abbotts. Nuclear power: All the king’s horses and all the king’s men …. Bulletin of the Atomic Scientists, 45(1):49– 52, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1990:PTR

[Abb90] John Abbotts. Perspective: Time for rebirth of civilian R&D. Bulletin of the Atomic Scientists, 46(9):12–13, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1991:BRN

[Abb91] John Abbotts. Book review: New Hampshire’s white elephant: Seabrook Station: Citizen Politics and Nuclear Power, by Henry F. Bedford. Bulletin of the Atomic Scientists, 47(2):42–43, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:1993:BRO

[Abb93] John Abbotts. Book review: Optimism by a nose: Life Under a Cloud: American Anxiety About the Atom, by Allan M. Winkler. Bulletin of the Atomic Scientists, 49(10):51–52, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi94a].

Abbas:1998:HC

[Abb98] Zaffar Abbas. The hardest choice. Bulletin of the Atomic Scientists, 54(4):36–37, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbotts:2004:HPI

[Abb04a] John Abbotts. Hanford: Piling it on. Bulletin of the Atomic Scientists, 60(3):??, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/3/23.

Abbotts:2004:LSD

[Abb04b] John Abbotts. The long, slow death of the Fast Flux Facility. Bulletin of the Atomic Scientists, 60(5):56–62, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/5/56.

Abbotts:2004:TP

[Abb04c] John Abbotts. The turning point. Bulletin of the Atomic Scientists, 60(5):66–68, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/5/66.

Abbott:2012:LGC

[Abb12] Derek Abbott. Limits to growth: Can nuclear power supply the world’s needs? Bulletin of the Atomic Scientists, 68(5):23–32, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/5/23.

Abelson:1974:SPG

[Abe74] Philip H. Abelson. A sport played by graduate students. Bulletin of the Atomic Scientists, 30(5):48–52, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:2002:LB

[ABH02] David Albright, Kathryn Buehler, and Holly Higgins. Bin laden and the bomb. Bulletin of the Atomic Scientists, 58(1):23–24, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/1/23.

Albright:2010:NKM

[ABKS10] David Albright, Paul Brannan, Robert Kelley, and Andrea Scheel Stricker. The North Korea–Myanmar relationship: a technical perspective. Bulletin of the Atomic Scientists, 66(3):19–29, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/3/19.

Able:1985:LDO

[Abl85] John Able. Letter: “Defense” and “offense” confused. Bulletin of the Atomic Scientists, 41(6):53, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gla85].

Abrahamson:1973:BBN

[Abr73] Dean E. Abrahamson. Books: The Nuclear-Power Rebellion: Citizens vs the Atomic Industrial Establishment, by Richard S. Lewis. Bulletin of the Atomic Scientists, 29(5):43–45, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrahamson:1974:BBL

[Abr74] Dean E. Abrahamson. Books: The Liquid Metal Fast Breeder Reactor: An Environmental and Economic Critique, by Thomas B. Cochran. Bulletin of the Atomic Scientists, 30(10):44–45, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrahamson:1976:LME

[Abr76] Dean E. Abrahamson. Letter: McCormack’s excesses. Bulletin of the Atomic Scientists, 32(3):4, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [O’L75].

Abrahamson:1979:GFC

[Abr79] Dean Abrahamson. Governments fall as consensus gives way to debate. Bulletin of the Atomic Scientists, 35(9):30–37, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1980:MCN

[Abr80] Herbert L. Abrams. Medical consequences of nuclear war. Bulletin of the Atomic Scientists, 36(7):15, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1983:CMC

[Abr83] Herbert L. Abrams. Civilian–military contingency hospital system: Preparing for “the highest rate of casualties in history”. Bulletin of the Atomic Scientists, 39(6):[various between 28–29], June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1984:MRA

[Abr84] Herbert L. Abrams. Medical resources after nuclear war. Bulletin of the Atomic Scientists, 40(10):23–26, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1986:CESa

[Abr86a] Herbert L. Abrams. Chernobyl: the emerging story: How radiation victims suffer. Bulletin of the Atomic Scientists, 42(7): 13–17, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1986:CESb

[Abr86b] Herbert L. Abrams. Chernobyl: the emerging story: The evacuation. Bulletin of the Atomic Scientists, 42(7):13, August/ September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1987:HIN

[Abr87] Herbert L. Abrams. Human instability and nuclear weapons. Bulletin of the Atomic Scientists, 43(1):34–39, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:2004:WMD

[Abr04] Herbert L. Abrams. Weapons of Miller’s descriptions. Bulletin of the Atomic Scientists, 60(4):56–64, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/56.

Abt:1984:BND

[Abt84] Clark C. Abt. Breaking the nuclear deadlock: Defensive role for the MX. Bulletin of the Atomic Scientists, 40(8):23–24, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amrine:1949:POS

[AC49] Michael E. Amrine and Edward A. Conway. The price of our survival. Bulletin of the Atomic Scientists, 5(11):300, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Audrieth:1965:SDS

[AC65] L. F. Audrieth and H. I. Chinn. The State Department seminar. Bulletin of the Atomic Scientists, 21(5):43–44, May 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adler:1976:LNS

[AC76] Howard I. Adler and David Dinsmore Comey. Letter: No suppression. Bulletin of the Atomic Scientists, 32(4):4, April 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com75a].

Albright:1985:FCN

[AC85] David Albright and André Carothers. Fragile consensus on Non-Proliferation Treaty. Bulletin of the Atomic Scientists, 41(11):8– 10, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Nat86b].

Auerbach:1986:LNF

[AC86] Arthur Auerbach and Malvine Cole. Letter: Nuclear fear and reality. Bulletin of the Atomic Scientists, 42(4):53–54, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SV86, Boy86a].

Abrahamson:2002:VAA

[AC02] James L. Abrahamson and Paul H. Carew. Vanguard of American atomic deterrence: the Sandia pioneers, 1946–1949. Praeger, Westport, CT, USA, 2002. ISBN 0-275-97819-2. xxiii + 181 pp. LCCN UG25 38th .A27 2002. URL http://www.loc.gov/

catdir/toc/fy035/2002070922.html. Foreword by Richard G.

Hewlett.

Alison:1967:CUS

[ACC^{+}67] Graham Alison, Jr., Steven L. Canby, Jack W. Carlson, John T. Dunlap, Charles Fried, Robert E. Herzstein, Samuel P. Huntington, Stephen A. Marglin, John Rawls, Gerald D. Rosenthal, Henry Rosovsky, Thomas C. Schelling, Lester C. Thurow, and Robert V. Zupkis. Comments: U.S. selective service. Bulletin of the Atomic Scientists, 23(8):38–40, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cro68] and reply [Sch68a].

Acheson:1956:NDA

[Ach56] Dean Acheson. A note from Dean Acheson. Bulletin of the Atomic Scientists, 12(4):134, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:RPU

[ACH84] William M. Arkin, Thomas B. Cochran, and Milton M. Hoenig. Resource paper on the U.S. nuclear arsenal. Bulletin of the Atomic Scientists, 40(7):1s–15s [various between 28–29], August/ September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Blu84a].

Acheson:2010:BNR

[Ach10] Ray Acheson. Beyond the 2010 NPT review conference: What’s next for nuclear disarmament? Bulletin of the Atomic Scientists, 66(6):77–87, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/66/6/77.

Acheson:2012:MNW

[Ach12] Ray Acheson. Modernization of nuclear weapons: Aspiring to “indefinite retention”? Bulletin of the Atomic Scientists, 68(5): 88–95, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/68/5/88.

Ackland:1984:SBT

[Ack84] Len Ackland. Scientists and the Bulletin tradition. Bulletin of the Atomic Scientists, 40(10):2, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1985:BRB

[Ack85] Len Ackland. Book review: Rethinking the Soviet Experience, by Stephen F. Cohen. Bulletin of the Atomic Scientists, 41(6): 52, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1986:PTW

[Ack86] Len Ackland. Perspective: Testing — who is cheating whom? Bulletin of the Atomic Scientists, 42(8):9–11, October 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:BRB

[Ack87a] Len Ackland. Book review: Nuclear Fallacy: Dispelling the Myth of Nuclear Strategy, by Morton H. Halperin. Bulletin of the Atomic Scientists, 43(6):54–55, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [You87a].

Ackland:1987:ETO

[Ack87b] Len Ackland. From the Editors: A time of opportunity. Bulletin of the Atomic Scientists, 43(2):2, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:ECG

[Ack87c] Len Ackland. From the Editors: Credibility gulf. Bulletin of the Atomic Scientists, 43(6):2, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:EMB

[Ack87d] Len Ackland. From the Editors: Missile begets new charge. Bulletin of the Atomic Scientists, 43(8):2, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:EMA

[Ack87e] Len Ackland. From the Editors: Much ado about zero. Bulletin of the Atomic Scientists, 43(7):2, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:EAT

[Ack87f] Len Ackland. From the Editors: The ABM Treaty and the Soviets. Bulletin of the Atomic Scientists, 43(3):2, April 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:EAD

[Ack87g] Len Ackland. From the Editors: The administration’s disarray. Bulletin of the Atomic Scientists, 43(1):2, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:EUT

[Ack87h] Len Ackland. From the Editors: Unconventional thinking. Bulletin of the Atomic Scientists, 43(10):2, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1987:EWH

[Ack87i] Len Ackland. From the Editors: Winds of hope. Bulletin of the Atomic Scientists, 43(9):2, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EIS

[Ack88a] Len Ackland. [Editorial introduction to Struggles of a Soviet scientist]. Bulletin of the Atomic Scientists, 44(8):33, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:ENS

[Ack88b] Len Ackland. [Editors note on the Soviet proposal for European security]. Bulletin of the Atomic Scientists, 44(7):8, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EPA

[Ack88c] Len Ackland. From the Editors: A “peace” attack? Bulletin of the Atomic Scientists, 44(6):2, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EHS

[Ack88d] Len Ackland. From the Editors: History set straight. Bulletin of the Atomic Scientists, 44(8):2, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EPW

[Ack88e] Len Ackland. From the Editors: Playing to win. Bulletin of the Atomic Scientists, 44(9):2, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EPI

[Ack88f] Len Ackland. From the Editors: Proliferation ignored. Bulletin of the Atomic Scientists, 44(4):2, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:ERP

[Ack88g] Len Ackland. From the Editors: Routine provocations. Bulletin of the Atomic Scientists, 44(3):2, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:ESM

[Ack88h] Len Ackland. From the Editors: Six minutes to midnight. Bulletin of the Atomic Scientists, 44(1):3, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [KA88, Hur88].

Ackland:1988:ESC

[Ack88i] Len Ackland. From the Editors: Spotlight on central Europe. Bulletin of the Atomic Scientists, 44(7):2, September 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:ETS

[Ack88j] Len Ackland. From the Editors: Taking stock. Bulletin of the Atomic Scientists, 44(2):2, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EPE

[Ack88k] Len Ackland. From the Editors: The perestroika experiment. Bulletin of the Atomic Scientists, 44(10):2, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:ETW

[Ack88l] Len Ackland. From the Editors: Third World missiles. Bulletin of the Atomic Scientists, 44(5):2, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:EUW

[Ack88m] Len Ackland. From the Editors: U.S. warhead production. Bulletin of the Atomic Scientists, 44(1):2, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1988:MWU

[Ack88n] Len Ackland. Making warheads: U.S. nuclear weapons production: an overview. Bulletin of the Atomic Scientists, 44(1):12, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pau88, Don88].

Ackland:1989:EAF

[Ack89a] Len Ackland. From the Editors: All the fiction fit to print — or broadcast. Bulletin of the Atomic Scientists, 45(9):2, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:EBE

[Ack89b] Len Ackland. From the Editors: Bush emerges. Bulletin of the Atomic Scientists, 45(6):2, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:ECW

[Ack89c] Len Ackland. From the Editors: Checking what he did. Bulletin of the Atomic Scientists, 45(1):2, January/February 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:ECE

[Ack89d] Len Ackland. From the Editors: Clear-eyed reassessments. Bulletin of the Atomic Scientists, 45(4):2, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:ECP

[Ack89e] Len Ackland. From the Editors: Clock power. Bulletin of the Atomic Scientists, 45(5):2, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:EOS

[Ack89f] Len Ackland. From the Editors: Open science, open skies. Bulletin of the Atomic Scientists, 45(7):2, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:EAS

[Ack89g] Len Ackland. From the Editors: The “atomic” scientists. Bulletin of the Atomic Scientists, 45(8):2, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1989:EUD

[Ack89h] Len Ackland. From the Editors: Unnatural disasters. Bulletin of the Atomic Scientists, 45(10):2, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pat90b].

Ackland:1990:ETM

[Ack90a] Len Ackland. Editorial: Ten minutes to midnight. Bulletin of the Atomic Scientists, 46(3):3, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:EAA

[Ack90b] Len Ackland. From the Editors: Atomic apartheid. Bulletin of the Atomic Scientists, 46(6):2, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:EBC

[Ack90c] Len Ackland. From the Editors: Bush critics miss the mark. Bulletin of the Atomic Scientists, 46(1):2, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:EDK

[Ack90d] Len Ackland. From the Editors: Democracy for Kuwait? Bulletin of the Atomic Scientists, 46(8):2, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:EHG

[Ack90e] Len Ackland. From the Editors: Human guinea pigs. Bulletin of the Atomic Scientists, 46(7):2, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:ELP

[Ack90f] Len Ackland. From the Editors: Lethal peace fallout. Bulletin of the Atomic Scientists, 46(4):2, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:ELA

[Ack90g] Len Ackland. From the Editors: Look ahead. Bulletin of the Atomic Scientists, 46(10):2, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:EMT

[Ack90h] Len Ackland. From the Editors: Messy tradeoffs. Bulletin of the Atomic Scientists, 46(5):2, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:EMA

[Ack90i] Len Ackland. From the Editors: Myopic arms deals. Bulletin of the Atomic Scientists, 46(9):2, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1990:ENP

[Ack90j] Len Ackland. From the Editors: Nuclear power revisited. Bulletin of the Atomic Scientists, 46(3):2, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [KA90].

Ackland:1990:EPD

[Ack90k] Len Ackland. From the Editors: Peacekeeper dividend. Bulletin of the Atomic Scientists, 46(2):2, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bon90].

Ackland:1991:EUW

[Ack91a] Len Ackland. From the Editors: An unnecessary war. Bulletin of the Atomic Scientists, 47(2):2, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wes91, Kle91].

Ackland:1991:EGM

[Ack91b] Len Ackland. From the Editors: Gulf macho. Bulletin of the Atomic Scientists, 47(1):2, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1991:EIA

[Ack91c] Len Ackland. From the Editors: Iraq’s apocalypse. Bulletin of the Atomic Scientists, 47(4):2, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1991:ENW

[Ack91d] Len Ackland. From the Editors: New world what? Bulletin of the Atomic Scientists, 47(5):2, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1991:EWM

[Ack91e] Len Ackland. From the Editors: The war metaphor. Bulletin of the Atomic Scientists, 47(3):2, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1992:WHW

[Ack92] Len Ackland. Who the hell will insure us? Bulletin of the Atomic Scientists, 48(9):22–27, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1994:RNW

[Ack94] Len Ackland. Report: Nuclear waste: A dump called Rocky Flats. Bulletin of the Atomic Scientists, 50(6):12–13, November/ December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1995:BRV

[Ack95] Len Ackland. Book review: Visions of victory: The hidden history of the Vietnam War, by John Prados. Bulletin of the Atomic Scientists, 51(6):66–68, November/December 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1997:BRC

[Ack97] Len Ackland. Book review: Camp cleanup: National Defense and the Environment, by Stephen Dycus, University Press of New England, 1996, 286 pages, $49.95 (cloth), $19.95 (paper). Bulletin of the Atomic Scientists, 53(2):58–60, March/April 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1999:DTA

[Ack99] Len Ackland. The day they almost lost Denver. Bulletin of the Atomic Scientists, 55(4):58–65, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/4/58.

Ackland:2001:RFC

[Ack01] Len Ackland. Rocky flats: Closing in on closure. Bulletin of the Atomic Scientists, 57(6):52–56, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/6/52.

Ackland:2007:RF

[Ack07] Len Ackland. Regional forecasts. Bulletin of the Atomic Scientists, 63(3):67–69, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/3/67.

Ackland:2009:CGS

[Ack09] Len Ackland. Can Germany survive without nuclear power? Bulletin of the Atomic Scientists, 65(4):41–52, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/4/41.

Acton:2009:MPR

[Act09] James M. Acton. The myth of proliferation-resistant technology. Bulletin of the Atomic Scientists, 65(6):49–59, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/6/49.

Allison:1954:SAF

[ACU^{+}54] Samuel K. Allison, Edward Condon, Harold C. Urey, Farrington Daniels, F. W. Loomis, Edward Shils, Linus Pauling, S. A. Goudsmit, Julian Schwinger, Albert Einstein, Hugh C. Wolfe, Cyril S. Smith, Victor F. Weisskopf, and Hans Bethe. Scientists affirm faith in Oppenheimer. Bulletin of the Atomic Scientists, 10(5):188–190, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AHCP:1990:OLP

[Ad 90] Ad Hoc Committee for Panama. Open letter to President George Bush [in protest of US invasion of Panama]. Bulletin of the Atomic Scientists, 46(2):50, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1947:RRP

[Ada47] Douglas P. Adams. Replies to Ridenour: A primary influence ignored. Bulletin of the Atomic Scientists, 3(8):229, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1952:BRB

[Ada52] Douglas P. Adams. Book review: An International Bibliography on Atomic Energy. Vol II. Scientific Aspects; Civil Defense in Modern War, by Augustin M. Prentiss. Bulletin of the Atomic Scientists, 8(5):164, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1961:BBP

[Ada61] Robert M. Adams. Books: Descriptive Palaeoclimatology, edited by A. E. M. Nairn. Bulletin of the Atomic Scientists, 17(10): 437, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1967:BBM

[Ada67] Robert McC. Adams. Books: Men, Machines, and Modern Times, by Elting E. Morison. Bulletin of the Atomic Scientists, 23(3):26, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1973:HAF

[Ada73] Franklin S. Adams. Hurricane Agnes: Flooding vs. dams in Pennsylvania. Bulletin of the Atomic Scientists, 29(4):30–34, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1982:WDW

[Ada82] Gordon Adams. What do weapons secure? Bulletin of the Atomic Scientists, 38(4):8–10, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1983:RDB

[Ada83] Gordon Adams. Reagan’s defense budget: Congress begins the debate. Bulletin of the Atomic Scientists, 39(4):25–28, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1986:ECM

[Ada86a] Gordon Adams. Economic conversion misses the point. Bulletin of the Atomic Scientists, 42(2):24–28, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bea86a, Ada86b].

Adams:1986:R

[Ada86b] Gordon Adams. A rejoinder. Bulletin of the Atomic Scientists, 42(6):50–51, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bea86a, Ada86a].

Adams:1986:WRP

[Ada86c] Gordon Adams. Washington report: Pentagon maneuvres for high funding. Bulletin of the Atomic Scientists, 42(4):5–7, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1987:BRB

[Ada87] Gordon Adams. Book reviews: The Defense Game, by Richard A. Stubbing with Richard A. Mendel; To Arm a Nation: Rebuilding America’s Endangered Defenses, by Richard Halloran; The New Maginot Line, by Jon Connell. Bulletin of the Atomic Scientists, 43(5):48, 50–51, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adamskii:1995:DMK

[Ada95] Viktor Adamskii. Dear Mr. Khrushchev. Bulletin of the Atomic Scientists, 51(6):28–31, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1997:RER

[Ada97] Ruth Adams. Recalling Eugene [Rabinowitch]. Bulletin of the Atomic Scientists, 53(6):45, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:2006:DR

[Ada06] Gordan Adams. Diminishing returns. Bulletin of the Atomic Scientists, 62(3):17–19, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/3/17.

Adedeji:1972:DI

[Ade72] Adebayo Adedeji. Deeds and intentions. Bulletin of the Atomic Scientists, 28(7):53, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adeniran:1982:BAR

[Ade82] Tunde Adeniran. Black Africa reacts. Bulletin of the Atomic Scientists, 38(7):36–38, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [NF83].

Adelman:1991:SJS

[Ade91] Kenneth L. Adelman. START: Just a sideshow. Bulletin of the Atomic Scientists, 47(9):19–21, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sha92a].

Adelman:2002:CC

[Ade02] Howard Adelman. Chaos in the camps. Bulletin of the Atomic Scientists, 58(6):88–93, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/6/88.

Anfinsen:1973:COL

[ADH^{+}73] Christian B. Anfinsen, Max Delbruck, H. K. Hartline, Arthur Kornberg, Julius Axelrod, Marshall Nirenberg, William H. Stein, Albert Szent-Györgyi, Edward L. Tatum, and Frank H. Westheimer. Communications: On open letter to President Nixon from the Federation of American Scientists, June 6, 1973. Bulletin of the Atomic Scientists, 29(7):2, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adie:1969:CSL

[Adi69] W. A. C. Adie. China’s “Second Liberation” in perspective. Bulletin of the Atomic Scientists, 25(2):12–16, 18–22, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Arkin:1991:NWH

[ADK91] William M. Arkin, Damian Durrant, and Hans Kristensen. Nuclear weapons headed for the trash. Bulletin of the Atomic Scientists, 47(10):14–19, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [You92].

Adler:1957:PVS

[Adl57] Mortimer Adler. Part VIII: Science and philosophy: The questions science cannot answer. Bulletin of the Atomic Scientists, 13(4):120–125, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adler:1969:BBS

[Adl69] Stephen L. Adler. Books: The Story of Quantum Mechanics, by Victor Guillemin. Bulletin of the Atomic Scientists, 25(8):41, 48, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adler:1985:SUS

[Adl85] Allan Adler. Secrecy: Unclassified secrets. Bulletin of the Atomic Scientists, 41(3):26–28, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adrian:1954:SHN

[Adr54] E. D. Adrian. Science and human nature. Bulletin of the Atomic Scientists, 10(9):338–341, 368, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ACNAS:1950:SFR

[Adv50] Advisory Committee of the National Academy of Sciences. Science and foreign relations: Berkner report to the U.S. Department of State. Bulletin of the Atomic Scientists, 6(10):293–297, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ACSTDAC:1960:DDS

[Adv60a] Advisory Committee on Science and Technology of the Democratic Advisory Council. Defense, disarmament, and survival. Bulletin of the Atomic Scientists, 16(4):137–138, 144, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ADSTDAC:1960:SPP

[Adv60b] Advisory Committee on Science and Technology of the Democratic Advisory Council. Science and party politics: Nuclear tests and national security. Bulletin of the Atomic Scientists, 16(5): 181–184, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ACUSPTIAEA:1962:UPT

[Adv62] Advisory Committee on U.S.Policy Toward the International Atomic Energy Agency. U.S. policy toward IAEA. Bulletin of the Atomic Scientists, 18(10):40–41, 43, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbess:1985:DRU

[AE85] Daniel Arbess and William Epstein. Disarmament role for the United Nations? Bulletin of the Atomic Scientists, 41(5):26–28, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aebersold:1948:ITA

[Aeb48] Paul Aebersold. Isotopes and their application to peacetime user of atomic energy. Bulletin of the Atomic Scientists, 4(5):151–154, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AIAG:1949:RAI

[AEC49] AEC Industrial Advisory Group. Report of the AEC Industrial Advisory Group. Bulletin of the Atomic Scientists, 5(2):51–55, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Lil49a].

Arkin:1985:NAF

[AF85] William M. Arkin and Richard W. Fieldhouse. Nuclear allergy: Focus on the nuclear infrastructure. Bulletin of the Atomic Scientists, 41(6):11–15, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1988:MNF

[AF88] David Albright and Martha Fell. A military nuclear fuel cycle primer. Bulletin of the Atomic Scientists, 44(1):15, January/ February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Afheldt:1988:NPO

[Afh88] Horst Afheldt. New policies, old fears. Bulletin of the Atomic Scientists, 44(7):24–28, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aftergood:1986:NSM

[Aft86] Steven Aftergood. Nuclear space mishaps and Star Wars. Bulletin of the Atomic Scientists, 42(8):40–43, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aftergood:1994:NLL

[Aft94a] Steven Aftergood. A non-lethal laundry list. Bulletin of the Atomic Scientists, 50(5):43, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aftergood:1994:SKF

[Aft94b] Steven Aftergood. The soft-kill fallacy. Bulletin of the Atomic Scientists, 50(5):40–45, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [ARA94a].

Aftergood:1999:WP

[Aft99] Steven Aftergood. Wrongheaded “Protection”. Bulletin of the Atomic Scientists, 55(4):6, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/4/6.

Aftergood:2000:SBF

[Aft00] Steven Aftergood. Secrecy is back in fashion. Bulletin of the Atomic Scientists, 56(6):24–30, November/December 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/6/24.

Aftergood:2001:HI

[Aft01] Steven Aftergood. Homer II. Bulletin of the Atomic Scientists, 57(2):71–73, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/2/71.

Abt:1963:LEA

[AG63] C. C. Abt and Todd Gitlin. Letter to the Editor: After Cuba. Bulletin of the Atomic Scientists, 19(6):31, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pan63, Rab63b].

Adams:1986:RMS

[AG86] Gordon Adams and David Gold. Recasting the military spending debate. Bulletin of the Atomic Scientists, 42(8):26–32, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [All86, Pri86b, Vog87].

Arndt:1991:LPN

[AG91] John Arndt and Carl A. Goldstein. Letters: Pollution and nuclear power. Bulletin of the Atomic Scientists, 47(7):47, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ham91b, Mye91].

Albright:1997:RPF

[AG97] David Albright and Corey Gay. Report: Proliferation: A flash from the past. Bulletin of the Atomic Scientists, 53(6):15–17, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:TNN

[AG98] David Albright and Corey Gay. Taiwan: Nuclear nightmare averted. Bulletin of the Atomic Scientists, 54(1):54–60, January/ February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AgaKhan:1979:NPD

[Aga79] Sadruddin Aga Khan. The nuclear power debate in Western Europe. Bulletin of the Atomic Scientists, 35(7):11–12, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abt:2022:WBW

[AGB22] Robert Abt, Christopher Galik, and Justin Baker. When burning wood to generate energy makes climate sense. Bulletin of the Atomic Scientists, 78(3):152–157, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2062941.

Alfonsin:1988:SDD

[AGC^{+}88] Raul Alfonsin, Rajiv Gandhi, Ingvar Carlsson, Andreas Papandreou, Miguel de la Madrid, and Julius Nyerere. Stockholm declaration on disarmament. Bulletin of the Atomic Scientists, 44(8): 44–45, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gol88e].

Agnew:1977:PLS

[Agn77a] Harold M. Agnew. A plan to lesson suspicions. Bulletin of the Atomic Scientists, 33(3):6–7, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ken77].

Agnew:1977:PER

[Agn77b] Harold M. Agnew. A primer on enhanced radiation weapons. Bulletin of the Atomic Scientists, 33(10):6–8, December 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mus78].

Agnew:1982:CLS

[Agn82a] Harold M. Agnew. Commentary: Leasing: a solution to nuclear proliferation. Bulletin of the Atomic Scientists, 38(2):40– 41, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Agnew:1982:EI

[Agn82b] Harold M. Agnew. Early impressions. Bulletin of the Atomic Scientists, 38(10):20–21, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano83j] that corrects the photo caption: it shows the first nuclear pile after reassembly at Argonne National Laboratory, not the original pile at Stagg Field.

Agosto:1977:LLA

[Ago77] William N. Agosto. Letter: Lucid argument. Bulletin of the Atomic Scientists, 33(7):2, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav77a].

Albright:1991:PRR

[AGZ91] David Albright, Peter Gray, and Tom Zamora. Perspective: Retire Rocky Flats. Bulletin of the Atomic Scientists, 47(10):12–13, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anspach:1962:LED

[AH62] Ralph Anspach and George E. Hlavka. Letter to the Editor: Disarmament. Bulletin of the Atomic Scientists, 18(7):36–37, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abbate:1979:EE

[AH79] Fred J. Abbate and Peter Harnik. Ethics and energy. Bulletin of the Atomic Scientists, 35(6):62–63, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Har79a].

Arkin:1989:AMN

[AH89a] William M. Arkin and Joshua M. Handler. Ask me no questions …. Bulletin of the Atomic Scientists, 45(6):24, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1989:NDS

[AH89b] William M. Arkin and Joshua M. Handler. Nuclear disasters at sea, then and now. Bulletin of the Atomic Scientists, 45(6):20–24, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:BHI

[AH91a] David Albright and Mark Hibbs. Bomb hype II. Bulletin of the Atomic Scientists, 47(7):16, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:HIB

[AH91b] David Albright and Mark Hibbs. Hyping the Iraqi bomb. Bulletin of the Atomic Scientists, 47(2):26–28, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:IBW

[AH91c] David Albright and Mark Hibbs. Iraq and the bomb: Were they even close? Bulletin of the Atomic Scientists, 47(2):16–25, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:INC

[AH91d] David Albright and Mark Hibbs. Iraq’s “nuclear complex”. Bulletin of the Atomic Scientists, 47(2):24, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:INH

[AH91e] David Albright and Mark Hibbs. Iraq’s nuclear hide-and-seek. Bulletin of the Atomic Scientists, 47(7):14–23, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:MRC

[AH91f] David Albright and Mark Hibbs. Making an running calutrons. Bulletin of the Atomic Scientists, 47(7):18–19, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:MBP

[AH91g] David Albright and Mark Hibbs. Making the bomb: producing fissile material. Bulletin of the Atomic Scientists, 47(2):20, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:MBE

[AH91h] David Albright and Mark Hibbs. Making the bomb: the enrichment plant. Bulletin of the Atomic Scientists, 47(2):21, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:MBG

[AH91i] David Albright and Mark Hibbs. Making the bomb: the gas centrifuge. Bulletin of the Atomic Scientists, 47(2):22, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:MBR

[AH91j] David Albright and Mark Hibbs. Making the bomb: the rotor assembly. Bulletin of the Atomic Scientists, 47(2):23, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:MBS

[AH91k] David Albright and Mark Hibbs. Making the bomb: the shopping list. Bulletin of the Atomic Scientists, 47(2):18, March 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:OP

[AH91l] David Albright and Mark Hibbs. Other paths. Bulletin of the Atomic Scientists, 47(7):21, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:RIN

[AH91m] David Albright and Mark Hibbs. Report: Iraq: News the front page missed. Bulletin of the Atomic Scientists, 47(8):7–9, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:RNP

[AH91n] David Albright and Mark Hibbs. Report: Nuclear proliferation: What’s North Korea up to, anyway? Bulletin of the Atomic Scientists, 47(10):10–11, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:DG

[AH92a] David Albright and Mark Hibbs. Digging for gold. Bulletin of the Atomic Scientists, 48(3):28, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:GCP

[AH92b] David Albright and Mark Hibbs. The gas centrifuge, piece by piece. Bulletin of the Atomic Scientists, 48(3):32–33, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:HHH

[AH92c] David Albright and Mark Hibbs. Hostages, headlines, and hype. Bulletin of the Atomic Scientists, 48(1):32, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:ISB

[AH92d] David Albright and Mark Hibbs. India’s silent bomb. Bulletin of the Atomic Scientists, 48(7):27–31, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Muk93].

Albright:1992:IBB

[AH92e] David Albright and Mark Hibbs. Iraq’s bomb: Blueprints and artifacts. Bulletin of the Atomic Scientists, 48(1):30–40, January/ February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [PB92].

Albright:1992:IST

[AH92f] David Albright and Mark Hibbs. Iraq’s shop-till-you-drop nuclear program. Bulletin of the Atomic Scientists, 48(3):27–37, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:LAK

[AH92g] David Albright and Mark Hibbs. The letters of Abdul Khan. Bulletin of the Atomic Scientists, 48(6):40–41, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:NKP

[AH92h] David Albright and Mark Hibbs. North Korea’s plutonium puzzle. Bulletin of the Atomic Scientists, 48(9):36–40, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:PBC

[AH92i] David Albright and Mark Hibbs. Pakistan’s bomb: out of the closet. Bulletin of the Atomic Scientists, 48(6):38–43, July/ August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:PP

[AH92j] David Albright and Mark Hibbs. Procurers and panderers. Bulletin of the Atomic Scientists, 48(1):34, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:RIA

[AH92k] David Albright and Mark Hibbs. Report: Iraq: It’s all over at Al Atheer. Bulletin of the Atomic Scientists, 48(5):8–10, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1992:RNP

[AH92l] David Albright and Mark Hibbs. Report: Nuclear proliferation: Spotlight shifts to Iran. Bulletin of the Atomic Scientists, 48(2): 9–11, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:RIS

[AH93a] David Albright and Mark Hibbs. Report: Iraq: Supplier-spotting. Bulletin of the Atomic Scientists, 49(1):8–9, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:SAA

[AH93b] David Albright and Mark Hibbs. South Africa: the ANC and the atom bomb. Bulletin of the Atomic Scientists, 49(3):32–37, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:2001:ABD

[AH01] David Albright and Corey Hinderstein. Algeria: Big deal in the desert? Bulletin of the Atomic Scientists, 57(3):45–52, May/ June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/57/3/45.

Albright:2002:NKT

[AH02] David Albright and Holly Higgins. North Korea: It’s taking too Long. Bulletin of the Atomic Scientists, 58(1):56–61, January/ February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/58/
1/56.

Albright:2003:BU

[AH03a] David Albright and Holly Higgins. A bomb for the Ummah. Bulletin of the Atomic Scientists, 59(2):49–55, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/2/49.

Albright:2003:IFF

[AH03b] David Albright and Corey Hinderstein. Iran: Furor over fuel. Bulletin of the Atomic Scientists, 59(3):12–15, May/June 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/3/12.

Albright:2003:IPR

[AH03c] David Albright and Corey Hinderstein. Iran, player or rogue? Bulletin of the Atomic Scientists, 59(5):52–58, September/ October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
5/52.

Albright:2004:CC

[AH04a] David Albright and Corey Hinderstein. The centrifuge connection. Bulletin of the Atomic Scientists, 60(2):61–66, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/2/61.

Albright:2004:ICS

[AH04b] David Albright and Corey Hinderstein. Iran: Countdown to showdown. Bulletin of the Atomic Scientists, 60(6):67–73, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/6/67.

Allison:2020:WAC

[AH20] David M. Allison and Stephen Herzog. “What about China?” and the threat to US–Russian nuclear arms control. Bulletin of the Atomic Scientists, 76(4):200–205, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1778370.

Asokan:2022:CCW

[AH22] Asha Asokan and Ira Helfand. Climate change and water scarcity will increase risk of nuclear catastrophe in South Asia. Bulletin of the Atomic Scientists, 78(4):214–217, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2087382.

Ahearne:1990:RHT

[Ahe90] John F. Ahearne. Radiation and health: Telling the public about risks. Bulletin of the Atomic Scientists, 46(7):37–39, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ahearne:1991:BRS

[Ahe91] John F. Ahearne. Book review: The science of regulation: The Fifth Branch: Science Advisors as Policy Makers, by Sheila Jasanoff. Bulletin of the Atomic Scientists, 47(2):44–45, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1990:SDM

[AHK90] William M. Arkin, Joshua Handler, and Hans Kristensen. Soviets disarm, mysteriously. Bulletin of the Atomic Scientists, 46(4):7, 54–55, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ahmed:1974:PPP

[Ahm74] W. Ahmed. Population policy and the peasant. Bulletin of the Atomic Scientists, 30(6):29–35, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ahmad:1980:STI

[Ahm80] Aqueil Ahmad. Science and technology in India. Bulletin of the Atomic Scientists, 36(9):38–41, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ahmedullah:1999:ISN

[Ahm99] Mohammed Ahmedullah. India: Sugarcoating nuclear power. Bulletin of the Atomic Scientists, 55(5):??, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/5/16.

Ahmedullah:2000:LEE

[Ahm00a] Mohammed Ahmedullah. Let ’em eat nukes. Bulletin of the Atomic Scientists, 56(5):52–57, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/52.

Ahmedullah:2000:PAC

[Ahm00b] Mohammed Ahmedullah. Pakistan: After the coup. Bulletin of the Atomic Scientists, 56(1):14–16, January/February 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/1/14.

Ahmedullah:2000:TMB

[Ahm00c] Mohammed Ahmedullah. Tibet: Marketing the Buddhist message. Bulletin of the Atomic Scientists, 56(2):13–16, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/2/13.

Ahmedullah:2003:ASU

[Ahm03a] Mohammed Ahmedullah. “Arms sales: The U.S.–French tug of war”. Bulletin of the Atomic Scientists, 59(5):14–16, September/ October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
5/14.

Ahmedullah:2003:BSS

[Ahm03b] Mohammed Ahmedullah. Burma: Not-so-selective sanctions. Bulletin of the Atomic Scientists, 59(6):12–14, November/ December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
6/12.

Ahmedullah:2004:IIM

[Ahm04] Mohammed Ahmedullah. International Islam: Mahathir says “Get Smart”. Bulletin of the Atomic Scientists, 60(2):14– 16, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/2/14.

Ahn:2010:NKM

[Ahn10] Mun Suk Ahn. The North Korea–Myanmar relationship: a historical perspective. Bulletin of the Atomic Scientists, 66(3):30–37, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/66/3/30.

al-Istrabadi:1998:LAI

[aI98] Feisal Amin al Istrabadi. Letter: Agony of the innocents. Bulletin of the Atomic Scientists, 54(5):3, 66, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LC98b].

Aiken:1961:CWL

[Aik61] Frank Aiken. Can we limit the nuclear club? Bulletin of the Atomic Scientists, 17(7):263–266, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Artsimovitch:1962:TDS

[AIL^{+}62] L. A. Artsimovitch, D. R. Inglis, R. S. Leghorn, A. Rich, U. V. Riznichenko, and P. E. Tamm. Test detection statement. Bulletin of the Atomic Scientists, 18(9):41, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AFA:1955:SHA

[Air55] Air Force Association. “Survival in the Hydrogen Age”. Bulletin of the Atomic Scientists, 11(1):29–34, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ajami:1978:HRW

[Aja78] Fouad Ajami. Human rights and world order politics. Bulletin of the Atomic Scientists, 34(8):8–9, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrahamson:1979:SRS

[AJSB79] Dean Abrahamson, Thomas B. Johansson, Peter Steen, and Wendy Barnaby. Special report: Sweden debates its peaceful atom. Bulletin of the Atomic Scientists, 35(9):29, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [RAB^{+}80].

Adler-Karlsson:1981:CTS

[AK81] Gunnar Adler-Karlsson. Commentary: A tragic statistical muddle. Bulletin of the Atomic Scientists, 37(7):60, August/ September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1993:AWD

[AK93] William M. Arkin and Hans Kristensen. From the Alice-in-Wonderland department. Bulletin of the Atomic Scientists, 49 (6):26, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1995:ICC

[AK95a] David Albright and Robert Kelley. Has Iraq come clean at last? Bulletin of the Atomic Scientists, 51(6):53–55, 60–64, November/ December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1995:MPM

[AK95b] David Albright and Robert Kelley. Massive programs, meager results. Bulletin of the Atomic Scientists, 51(6):56–60, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:DUS

[AK98a] Joseph Albright and Marcia Kunstel. “Did the United States have any secrets?”. Bulletin of the Atomic Scientists, 54(1):50, 53, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:MS

[AK98b] Joseph Albright and Marcia Kunstel. “Mlad” and “Star”. Bulletin of the Atomic Scientists, 54(1):48–52, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:DD

[AK98c] William M. Arkin and Hans Kristensen. Dangerous directions. Bulletin of the Atomic Scientists, 54(2):26–31, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Averre:2001:CWD

[AK01] Derek Averre and Igor Khripunov. Chemical weapons disposal: Russia tries again. Bulletin of the Atomic Scientists, 57(5):57– 63, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/5/57.

Albright:2004:FMS

[AK04] David Albright and Kimberly Kramer. Fissile material: Stockpiles still growing. Bulletin of the Atomic Scientists, 60(6):14–16, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/6/14.

Akhtar:2017:MNR

[Akh17] Rabia Akhtar. Managing nuclear risk in South Asia. Bulletin of the Atomic Scientists, 73(1):62–63, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2016.
1264217.

Akiba:2005:PS

[Aki05] Tadatoshi Akiba. Postscript: The survivors. Bulletin of the Atomic Scientists, 61(4):63, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AlKaabi:2011:UAE

[Al 11] Hamad Al Kaabi. United Arab Emirates and the experience of a nuclear newcomer. Bulletin of the Atomic Scientists, 67(4): 52–59, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/4/52.

AlAssad:2015:ZRT

[Al 15] Wael Al Assad. Is “zero” the right target for disarmament?: an Arab response. Bulletin of the Atomic Scientists, 71(1):91–94, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1962:LEM

[AL62] D. R. Anderson and R. P. Loomba. Letter to the Editor: Monitoring basic research. Bulletin of the Atomic Scientists, 18(8): 32, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Angier:1979:LPB

[AL79a] John Angier and Sanford A. Lakoff. Letter: PBS’ Nova. Bulletin of the Atomic Scientists, 35(3):87, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lak78].

Ash:1979:CFI

[AL79b] Arlene Ash and John Lamperti. Commentary: Funding for intervenors. Bulletin of the Atomic Scientists, 35(9):10–11, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albrecht:1954:LEM

[Alb54] Herbert O. Albrecht. Letter to the Editor: Madame Joliot and the American Chemical Society. Bulletin of the Atomic Scientists, 10(6):212, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albrecht:1957:LES

[Alb57] Herbert O. Albrecht. Letter to the Editor: Subjects for articles. Bulletin of the Atomic Scientists, 13(8):313, October 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albrecht:1961:WE

[Alb61] William A. Albrecht. Wastebasket of the Earth. Bulletin of the Atomic Scientists, 17(8):335–340, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albrecht:1964:LEF

[Alb64] Herbert O. Albrecht. Letter to the Editor: Fluoridation once again. Bulletin of the Atomic Scientists, 20(8):38, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar64h].

Albrecht:1971:CRL

[Alb71] Herbert O. Albrecht. Communications: Radiation limits. Bulletin of the Atomic Scientists, 27(3):2, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bou70].

Albers:1973:SMR

[Alb73] John P. Albers. Seabed mineral resources: A survey. Bulletin of the Atomic Scientists, 29(8):33–38, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albrecht:1975:LCC

[Alb75] Herbert O. Albrecht. Letters: Causes for concern. Bulletin of the Atomic Scientists, 31(8):3, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys75].

Albertini:1978:CAH

[Alb78] James V. Albertini. Commentary: Action in Hawaii. Bulletin of the Atomic Scientists, 34(10):57, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albertini:1981:SRH

[Alb81] James V. Albertini. Special report: Hawaii: life under the gun. Bulletin of the Atomic Scientists, 37(3):50–57, March 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [FBH81].

Albrecht:1983:EST

[Alb83] Ulrich Albrecht. European security through political conciliation. Bulletin of the Atomic Scientists, 39(5):50–52, May 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1984:FMP

[Alb84] David Albright. French military plans for Superphénix? Bulletin of the Atomic Scientists, 40(9):30–34, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [HA85].

Albright:1986:CUN

[Alb86] David Albright. Chernobyl and the U.S. nuclear industry. Bulletin of the Atomic Scientists, 42(9):38–40, November 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albrecht:1987:IVR

[Alb87a] Ulrich Albrecht. International views: Revive the Reykjavik dynamism. Bulletin of the Atomic Scientists, 43(2):40–41, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1987:BTB

[Alb87b] David Albright. Bombs without test blasts? Bulletin of the Atomic Scientists, 43(5):32, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1987:PBM

[Alb87c] David Albright. Pakistan’s bomb-making capacity. Bulletin of the Atomic Scientists, 43(5):30–33, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1989:BPS

[Alb89a] David Albright. Bomb potential for South America. Bulletin of the Atomic Scientists, 45(4):16–20, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1989:HUS

[Alb89b] David Albright. Hibbs uncovers scandals. Bulletin of the Atomic Scientists, 45(5):22, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1989:HDA

[Alb89c] David Albright. How did Argentina do it? Bulletin of the Atomic Scientists, 45(4):19, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1989:HDB

[Alb89d] David Albright. How did Brazil do it? Bulletin of the Atomic Scientists, 45(4):20, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:BSS

[Alb93a] David Albright. Bulletin: Slow but steady. Bulletin of the Atomic Scientists, 49(6):5–6, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:BSA

[Alb93b] David Albright. Bulletin: South Africa comes clean. Bulletin of the Atomic Scientists, 49(4):3–5, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:EH

[Alb93c] David Albright. Engineer for hire. Bulletin of the Atomic Scientists, 49(10):28–36, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:PP

[Alb93d] David Albright. A proliferation primer. Bulletin of the Atomic Scientists, 49(5):14–23, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1993:RNP

[Alb93e] David Albright. Report: Non-proliferation treaty: North Korea drops out. Bulletin of the Atomic Scientists, 49(4):9–11, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [FM93].

Albright:1993:RSA

[Alb93f] David Albright. Report: South Africa: A curious conversion. Bulletin of the Atomic Scientists, 49(5):8–11, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:BEC

[Alb94a] David Albright. Beating export controls. Bulletin of the Atomic Scientists, 50(4):41, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:BNK

[Alb94b] David Albright. Bulletin: North Korea and the “worst-case” scare-nario. Bulletin of the Atomic Scientists, 50(1):3–6, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:FA

[Alb94c] David Albright. The flash in the Atlantic. Bulletin of the Atomic Scientists, 50(4):42, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:HMP

[Alb94d] David Albright. How much plutonium does North Korea have? Bulletin of the Atomic Scientists, 50(5):46–53, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:RNS

[Alb94e] David Albright. Rust never sleeps. Bulletin of the Atomic Scientists, 50(5):49, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:SAA

[Alb94f] David Albright. South Africa and the affordable bomb. Bulletin of the Atomic Scientists, 50(4):37–47, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1994:UTU

[Alb94g] David Albright. Uranium tucked under the mattress? Bulletin of the Atomic Scientists, 50(4):38, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1995:IB

[Alb95] David Albright. An Iranian bomb? Bulletin of the Atomic Scientists, 51(4):20–26, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:POS

[Alb98a] David Albright. Pakistan: The other shoe drops. Bulletin of the Atomic Scientists, 54(4):24–25, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:RHD

[Alb98b] David Albright. Reports of his death were greatly exaggerated. Bulletin of the Atomic Scientists, 54(5):30–31, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:SHR

[Alb98c] David Albright. The shots heard ’round the world: India conducted three nuclear tests on May 11 and two on May 13. Bulletin of the Atomic Scientists, 54(4):20–25, July/August 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1998:TGM

[Alb98d] David Albright. Trouble in the Gulf: Masters of deception. Bulletin of the Atomic Scientists, 54(3):44–50, May/June 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:2000:SM

[Alb00] David Albright. Secrets that matter. Bulletin of the Atomic Scientists, 56(6):57–62, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/6/57.

Albright:2006:WCI

[Alb06] David Albright. When could Iran get the bomb? Bulletin of the Atomic Scientists, 62(4):26–33, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/4/26.

Alcock:1963:BBO

[Alc63] Norman Alcock. Books: Peace and Opinion, by Evan Luard; The Price of Peace, by James J. Wadsworth. Bulletin of the Atomic Scientists, 19(1):31–32, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alcalay:1981:CPL

[Alc81] Glenn H. Alcalay. Commentary: On ‘Paradise Lost’. Bulletin of the Atomic Scientists, 37(5):63, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Joh80b].

Aldridge:1976:HDI

[Ald76] Robert C. Aldridge. How defense industries keep the business coming. Bulletin of the Atomic Scientists, 32(5):44–46, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aldy:2017:RWH

[Ald17] Joseph E. Aldy. Real world headwinds for Trump climate change policy. Bulletin of the Atomic Scientists, 73(6):376–381, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1388673.

Alexander:1960:LEW

[Ale60] William M. Alexander. Letter to the Editor: The weapons of suicide. Bulletin of the Atomic Scientists, 16(10):423, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alexander:1964:PO

[Ale64] Archibald S. Alexander. The problem and the opportunity. Bulletin of the Atomic Scientists, 20(4):8–10, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alexandrov:1967:HD

[Ale67] A. P. Alexandrov. The heroic deed. Bulletin of the Atomic Scientists, 23(10):10–11, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alexander:1969:BRP

[Ale69] Lyle T. Alexander. Book review: The price of pure water: Desalination: Water for Mankind’s Future, by Roy Popkin. Bulletin of the Atomic Scientists, 25(6):41, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alfven:1972:EE

[Alf72] Hannes Alfven. Energy and environment. Bulletin of the Atomic Scientists, 28(5):5–7, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mer72].

Alfven:1974:FEO

[Alf74] Hannes Alfven. Fission energy and other sources of energy. Bulletin of the Atomic Scientists, 30(1):4–8, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alfven:1977:C

[Alf77] Kerstin Alfven. A conversation. Bulletin of the Atomic Scientists, 33(5):43–44, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alfven:1979:SPD

[Alf79] Hannes Alfvén. Science, progress and destruction. Bulletin of the Atomic Scientists, 35(3):68–71, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alfven:1981:HIV

[Alf81] Hannes Alfvén. Human IQ vs. nuclear IQ. Bulletin of the Atomic Scientists, 37(1):4–5, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mal81].

Alford:1984:BRB

[Alf84] Jonathan Alford. Book review: Conventional Deterrence, by John J. Mearsheimer. Bulletin of the Atomic Scientists, 40(5):41–42, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alison:1957:BPU

[Ali57] Helen C. Alison. British Parliament and U.S. Secretary of State look at arms policy. Bulletin of the Atomic Scientists, 13(9):336– 340, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1948:DAS

[All48] Samuel K. Allison. Dr. Allison speaks in support of Dr. Condon. Bulletin of the Atomic Scientists, 4(4):98, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1950:SPP

[All50] Samuel K. Allison. The state of physics; or the perils of being important. Bulletin of the Atomic Scientists, 6(1):2–4, 26–27, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1951:RUP

[All51] Samuel K. Allison. The responsibility of a university professor. Bulletin of the Atomic Scientists, 7(12):367–370, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1953:TYA

[All53] Samuel K. Allison. Ten years of the Atomic Age. Bulletin of the Atomic Scientists, 9(1):8–10, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1954:BBA

[All54a] Helen C. Allison. Books: Britain’s Atomic Factories, by K. E. B. Jay. Bulletin of the Atomic Scientists, 10(10):394, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1954:BBC

[All54b] Helen C. Allison. Books: Removing the Causes of War, by Kathleen Lonsdale. Bulletin of the Atomic Scientists, 10(8):330, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1954:NR

[All54c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 10(10):395–397, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1954:BBN

[All54d] Samuel K. Allison. Books: Introduction to Nuclear Engineering, by Richard Stephenson. Bulletin of the Atomic Scientists, 10 (6):233, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:BRBc

[All55a] Helen C. Allison. Book review: World Development of Atomic Energy, by Atomic Industry Forum, Inc. Edited by Oliver Townsend. Bulletin of the Atomic Scientists, 11(9):349, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:BRBb

[All55b] Helen C. Allison. Book review: World Federal Government, Why? What? How? In Maximum Terms, by Erdith Wynner. Bulletin of the Atomic Scientists, 11(3):95, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRa

[All55c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(1):38–40, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRb

[All55d] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(2):69–72, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRc

[All55e] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(3):102–103, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRe

[All55f] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(5):198–199, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRf

[All55g] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(6):231–232, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRg

[All55h] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(7):269–270, 273, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRj

[All55i] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(8):311–313, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRk

[All55j] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(9):350–352, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:NRl

[All55k] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 11(10):377–378, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:BRBa

[All55l] Samuel K. Allison. Book review: High Energy Accelerators, by M. Stanley Livingston. Bulletin of the Atomic Scientists, 11(3):94, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRa

[All56a] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(1):29–31, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRb

[All56b] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(2):62–64, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRc

[All56c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(3):94–96, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRd

[All56d] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(4):136–137, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRe

[All56e] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(5):181–183, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRf

[All56f] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(6):229–231, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRg

[All56g] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(7):273–275, 279, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRh

[All56h] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(8):318–319, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRi

[All56i] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(9):350–352, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:NRj

[All56j] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 12(10):378–380, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:BBH

[All56k] Samuel K. Allison. Books: Atom Harvest, by Leonard Bertin. Bulletin of the Atomic Scientists, 12(6):231, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1956:BBB

[All56l] Samuel K. Allison. Books: Niels Bohr and the Development of Physics, edited by W. Pauli with the assistance of L. Rosenfeld and V. Weisskopf. Bulletin of the Atomic Scientists, 12(2):61, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRa

[All57a] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(1):36–38, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRb

[All57b] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(2):68–70, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRc

[All57c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(3):110–112, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRd

[All57d] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(4):147–149, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRe

[All57e] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(5):181–183, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRf

[All57f] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(6):229–231, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRg

[All57g] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(7):267–271, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRh

[All57h] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(8):309–312, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRi

[All57i] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(9):343–344, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1957:NRj

[All57j] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 13(10):373–374, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alloy:1957:LEM

[All57k] Evelyn Alloy. Letter to the Editor: The manufacture and testing of atomic weapons. Bulletin of the Atomic Scientists, 13(6):232, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:JMN

[All58a] Helen C. Allison. John McCone: New AEC Chairman. Bulletin of the Atomic Scientists, 14(8):334–335, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRa

[All58b] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(1):62–64, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRb

[All58c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(2):94–96, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRc

[All58d] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(3):126–128, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRd

[All58e] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(4):157–159, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRe

[All58f] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(5):196–199, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRf

[All58g] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(6):238–240, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRg

[All58h] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(7):278–280, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRh

[All58i] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(8):331–333, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRi

[All58j] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(9):396–399, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1958:NRj

[All58k] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 14(10):429–430, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:BDU

[All59a] Helen C. Allison. Bethe on detection of underground explosions. Bulletin of the Atomic Scientists, 15(6):257, 259, June 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRa

[All59b] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(1):47–48, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRb

[All59c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(2):94–97, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRc

[All59d] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(3):144–145, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRd

[All59e] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(4):174–176, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRe

[All59f] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(5):223–224, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRf

[All59g] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(6):269–270, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRg

[All59h] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(7):317–320, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRh

[All59i] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(8):351–353, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRi

[All59j] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(9):398–400, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:NRj

[All59k] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 15(10):428–430, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1959:BBE

[All59l] Samuel K. Allison. Books: Magnets: The Education of a Physicist, by Francis Bitter; The Neutron Story, by Donald J. Hughes; Soap Bubbles and the Forces Which [sic] Mould Them, by C. V. Boys. Bulletin of the Atomic Scientists, 15(8):350, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:NRa

[All60a] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 16(1):30–32, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:NRb

[All60b] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 16(2):78–80, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:NRf

[All60c] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 16(6):223–225, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:NRg

[All60d] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 16(7):302–305, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:NRh

[All60e] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 16(8):342–344, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:NRi

[All60f] Helen C. Allison. News roundup. Bulletin of the Atomic Scientists, 16(9):375–376, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1960:OSL

[All60g] Helen C. Allison. Outspoken scientist: Linus Carl Pauling. Bulletin of the Atomic Scientists, 16(10):382, 390, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano61h] and comments [Rab61j].

Allison:1960:PEN

[All60h] Samuel K. Allison. Physics in Egypt: A new type of lend-lease. Bulletin of the Atomic Scientists, 16(8):317–321, 335, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1962:PS

[All62a] Helen C. Allison. Protein from the sea. Bulletin of the Atomic Scientists, 18(7):42–43, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1962:SC

[All62b] Helen C. Allison. Scientist as citizen. Bulletin of the Atomic Scientists, 18(5):34–35, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allen:1965:LED

[All65] Daniel Allen. Letter to the Editor: Deterring the ‘deterers’. Bulletin of the Atomic Scientists, 21(8):29–30, October 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tal65].

Alland:1968:WDA

[All68] Alexander Alland, Jr. War and disease: An anthropological perspective. Bulletin of the Atomic Scientists, 24(6):28–31, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allain:1986:LEM

[All86] R. Jay Allain. Letter: Economics and morality. Bulletin of the Atomic Scientists, 42(10):51, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AG86].

Allain:1988:LTE

[All88] R. Jay Allain. Letter: Tangled ethics. Bulletin of the Atomic Scientists, 44(6):53–54, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor88].

Allison:2006:OFI

[All06] Graham Allison. The ongoing failure of imagination. Bulletin of the Atomic Scientists, 62(5):34–41, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/5/34.

Allenby:2014:NTU

[All14a] Braden R. Allenby. Are new technologies undermining the laws of war? Bulletin of the Atomic Scientists, 70(1):21–31, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/70/
1/21.full.pdf+html.

Allenby:2014:SST

[All14b] Braden R. Allenby. Senior soldiers: The thin gray line. Bulletin of the Atomic Scientists, 70(5):86–95, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allenby:2015:ETF

[All15a] Brad Allenby. Emerging technologies and the future of humanity. Bulletin of the Atomic Scientists, 71(6):29–38, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allenby:2015:PDA

[All15b] Braden R. Allenby. The paradox of dominance: The age of civilizational conflict. Bulletin of the Atomic Scientists, 71(2):60–74, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allenby:2016:FSO

[All16] Brad Allenby. Frankensteins and space odysseys: Our history with technology, our future with machines. Bulletin of the Atomic Scientists, 72(4):256–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allenby:2017:HDD

[All17] Brad Allenby. Here be dragons: DARPA in the age of hybrid war. Bulletin of the Atomic Scientists, 73(3):188–191, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1315041.

Allenby:2018:DWA

[All18] Brad Allenby. Designer warriors: Altering conflict — and humanity itself? Bulletin of the Atomic Scientists, 74(6):379–384, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1533194.

Almond:1985:NWL

[Alm85] Harry H. Almond, Jr. Nuclear weapons are legal tools. Bulletin of the Atomic Scientists, 41(5):32–35, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alperin:1962:PPI

[Alp62] Robert J. Alperin. The public, policy, and information. Bulletin of the Atomic Scientists, 18(8):20–22, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alperovitz:1985:MAD

[Alp85] Gar Alperovitz. More on atomic diplomacy. Bulletin of the Atomic Scientists, 41(11):35–39, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alperovitz:1992:AQS

[Alp92] Gar Alperovitz. Author’s query: [study on the bombing of Hiroshima]. Bulletin of the Atomic Scientists, 48(3):46, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alsop:1951:BB

[Als51] Stewart Alsop. The British and the bomb. Bulletin of the Atomic Scientists, 7(6):164, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1974:BLL

[Alv74] Luis W. Alvarez. Berkeley: A lab like no other. Bulletin of the Atomic Scientists, 30(4):18–23, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1980:CRE

[Alv80] Robert Alvarez. Commentary: Radiation exposure limits. Bulletin of the Atomic Scientists, 36(9):58–59, November 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1984:RSB

[Alv84] Robert Alvarez. Radiation standards and A-bomb survivors. Bulletin of the Atomic Scientists, 40(8):26–28, October 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1987:BRB

[Alv87] Robert Alvarez. Book review: The Dragon’s Tail: Radiation Safety in the Manhattan Project, 1942–1946, by Barton Hacker. Bulletin of the Atomic Scientists, 43(10):46, 48, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:2000:ED

[Alv00] Robert Alvarez. Energy in decay. Bulletin of the Atomic Scientists, 56(3):24–35, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/3/24.

Alvarez:2001:LSD

[Alv01a] Robert Alvarez. A Long season of discontent. Bulletin of the Atomic Scientists, 57(1):52–57, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/52.

Alvarez:2001:MIW

[Alv01b] Robert Alvarez. Making IT work. Bulletin of the Atomic Scientists, 57(4):55–60, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/4/55.

Alvarez:2002:WAS

[Alv02] Robert Alvarez. What about the spent fuel? Bulletin of the Atomic Scientists, 58(1):45–47, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/45.

Alvarez:2003:NKN

[Alv03] Robert Alvarez. North Korea: No bygones at Yongbyon. Bulletin of the Atomic Scientists, 59(4):38–45, July/August 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/4/38.

Alvarez:2005:NWC

[Alv05] Robert Alvarez. Nuclear waste: Cleanup curveball. Bulletin of the Atomic Scientists, 61(4):16–18, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/4/16.

Alvarez:2014:NWD

[Alv14] Robert Alvarez. The nuclear weapons dismantlement problem. Bulletin of the Atomic Scientists, 70(6):22–28, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:2020:NKN

[Alv20] Robert Alvarez. 2000: North Korea: No bygones at Yongbyon. Bulletin of the Atomic Scientists, 76(6):404–409, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847509.

Altmann:1951:DSR

[AM51] Berthold Altmann and Harry Moskowitz. Dispersal — a selected reading list. Bulletin of the Atomic Scientists, 7(9):280–284, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Appel:1974:HML

[AM74] John Appel and James J. MacKenzie. How much light do we really need?: Industry’s illumination standards help sell energy and light bulbs, but they are open to scientific challenge. Bulletin of the Atomic Scientists, 30(10):18–24, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ackland:1986:ANA

[AM86a] Len Ackland and Steven McGuire, editors. Assessing the nuclear age: selections from the Bulletin of the Atomic Scientists. Educational Foundation for Nuclear Science, Chicago, IL, USA, 1986. ISBN 0-941682-07-2, 0-941682-08-0 (paperback). xvii + 382 pp. LCCN U264 .A79 1986. US$12.95.

Arbess:1986:BRB

[AM86b] Daniel Arbess and Alex Miller. Book reviews: Strategic command and control , by Bruce Blair; The Button, by Daniel Ford; The Command and Control of Nuclear Forces, by Paul Bracken. Bulletin of the Atomic Scientists, 42(4):50–52, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:2021:GMB

[AM21] Robert Alvarez and Joseph Mangano. I gave my baby tooth to science: Project Sunshine’s role in the Limited Test Ban Treaty and cutting-edge pollution research. Bulletin of the Atomic Scientists, 77(6):312–317, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1989207.

Amati:1977:LPO

[Ama77] Daniele Amati. Letter: Proving one’s innocence. Bulletin of the Atomic Scientists, 33(8):3, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev77a].

Amacher:2003:YYO

[Ama03] Peter Amacher. You’re on your own-again. Bulletin of the Atomic Scientists, 59(3):34–43, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/59/3/34.

Amber:2001:NZB

[Amb01] Dave Amber. New Zealand: Blocking the way. Bulletin of the Atomic Scientists, 57(3):18–19, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/18.

Amdur:1965:LET

[Amd65] Irving Amdur. Letter to the Editor: Two strategies for peace. Bulletin of the Atomic Scientists, 21(8):31, October 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AASW:1947:MBW

[Ame47] American Association of Scientific Workers. Memorandum to the UN — on bacterial warfare. Bulletin of the Atomic Scientists, 3 (12):363–365, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allen:1950:TIS

[AME50] J. Garrott Allen, Peter V. Moulder, and Daniel M. Enerson. The treatment of irradiation sickness. Bulletin of the Atomic Scientists, 6(8–9):263–264, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

AAUP:1953:FU

[Ame53] American Association of University Professors. The freedom of universities. Bulletin of the Atomic Scientists, 9(5):191–192, 194, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amery:1966:LEP

[Ame66] H. Amery. Letter to the Editor: In politics. Bulletin of the Atomic Scientists, 22(4):31–32, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor66a].

Amery:1968:BSA

[Ame68] Al Amery. The Bulletin: A spokesman for all seasons? Bulletin of the Atomic Scientists, 24(7):40–41, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amery:1969:CCP

[Ame69] Al Amery. Communications: Classes and parties. Bulletin of the Atomic Scientists, 25(8):48, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab69k, SS69].

Amer:1980:CPB

[Ame80] Nabil M. Amer. Commentary: Personal boycott. Bulletin of the Atomic Scientists, 36(7):58–59, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amerisov:1986:CES

[Ame86] Alexander Amerisov. Chernobyl: the emerging story: A chronology of Soviet media coverage. Bulletin of the Atomic Scientists, 42(7):38–39, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amrine:1951:BRB

[Amr51] Michael Amrine. Book review: We of Nagasaki, by Takash Nagai. Bulletin of the Atomic Scientists, 7(5):149–153, May 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amrine:1954:OSH

[Amr54] Michael Amrine. Our submerged hope for a brighter world. Bulletin of the Atomic Scientists, 10(5):156–158, 192, May 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amster:1956:DD

[Ams56] Warren Amster. Design for deterrence. Bulletin of the Atomic Scientists, 12(5):164–165, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amsden:1976:LPP

[Ams76] Ted Amsden. Letter: Pyramid of peace. Bulletin of the Atomic Scientists, 32(5):4, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Andrews:1954:TAC

[And54] Stanley Andrews. The technical assistance concept. Bulletin of the Atomic Scientists, 10(4):126, 128, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1955:BRB

[And55] Herbert L. Anderson. Book review: The Accident, by Dexter Masters. Bulletin of the Atomic Scientists, 11(10):376, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1956:CWH

[And56] Clinton Anderson. Can we halt the race for atomic arms? Bulletin of the Atomic Scientists, 12(6):223–226, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1958:LEA

[And58] Earl T. Anderson. Letter to the Editor: Astronautics and the future. Bulletin of the Atomic Scientists, 14(6):234, June 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds58].

Anderson:1960:WTS

[And60] Senator Clinton Anderson. Why tests should or should not continue. Bulletin of the Atomic Scientists, 16(3):86–87, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1965:BBJ

[And65] Clinton P. Anderson. Books: The Journals of David E. Lilienthal, Volume I: The TVA Years, 1939–1945. Volume II: The Atomic Energy Years, 1945–1950; Nuclear Energy, Public Policy and the Law, edited by Edward J. Bloustein. Bulletin of the Atomic Scientists, 21(3):42–44, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1968:BBB

[And68] C. Arnold Anderson. Books: Universities: British, Indian, African, by Eric Ashby. Bulletin of the Atomic Scientists, 24(1): 50–51, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1971:BBS

[And71] Herbert L. Anderson. Books: The Scorpion and the Tarantula: The Struggle to Control Atomic Energy 1945–1949, by Joseph I. Lieberman. Bulletin of the Atomic Scientists, 27(5):47, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anders:1972:RRE

[And72] Edward Anders. Research review: Edward Anders on meteorites and the Moon. Bulletin of the Atomic Scientists, 28(10):32–38, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1973:EDC

[And73] Herbert L. Anderson. Early days of the chain reaction. Bulletin of the Atomic Scientists, 29(4):8–12, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1974:AOTb

[And74a] Herbert L. Anderson. “All in our time”: Fermi, Szilard, and Trinity. Bulletin of the Atomic Scientists, 30(8):40–47, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also part 1 [And74b]. See also comment [Kat75].

Anderson:1974:AOTa

[And74b] Herbert L. Anderson. “All in our time”: The legacy of Fermi and Szilard. Bulletin of the Atomic Scientists, 30(7):56–62, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Continued in [And74a]. See also comment [Kat75].

Anderson:1978:BRB

[And78a] Duncan Anderson. Book reviews: The Zapping of America: Microwaves, Their Deadly Risk, and the Coverup, by Paul Brodeur. Bulletin of the Atomic Scientists, 34(9):49, November 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1978:LAM

[And78b] Sue Anderson. Letter: Another mother for peace. Bulletin of the Atomic Scientists, 34(5):6–7, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:1981:FSM

[And81] J. Edward Anderson. First strike: myth or reality. Bulletin of the Atomic Scientists, 37(9):6–11, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SA82].

Andersen:1983:BRB

[And83] Robert Andersen. Book review: The Islamic Bomb, by Steve Weissman and Herbert Krosney. Bulletin of the Atomic Scientists, 39(3):42, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Andrikos:1985:NAB

[And85] Nikos Andrikos. Nuclear allergy: A Balkan nuclear-weapons-free zone. Bulletin of the Atomic Scientists, 41(6):29–31, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Andreyev:1993:NDK

[And93] Nikolai Andreyev. From “nyet” to “don’t know”. Bulletin of the Atomic Scientists, 49(1):21–25, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Andreev:1994:KGD

[And94] Nikolai Andreev. Kirgizstan: Grappling with democracy. Bulletin of the Atomic Scientists, 50(1):52–55, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anderson:2010:BAS

[And10] Marvin Anderson. Bulletin of the Atomic Scientists. The New York Review of Magazines, ??(??):??, ???? 2010. URL http:

//archives.jrn.columbia.edu/2010/nyrm.org/2010/05/12/
bulletin-of-the-atomic-scientists/index.html.

Anger:1960:VRB

[Ang60] Herbert S. Anger. Viruses and Rickettsiae as BW agents. Bulletin of the Atomic Scientists, 16(6):246–247, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Angell:1978:PNP

[Ang78] Norman Angell. Perspective: Nobel Peace Laureates: 1933. Bulletin of the Atomic Scientists, 34(10):1, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 12 June 1935, Oslo, Norway.

Anichkina:2015:LMD

[Ani15] Tatiana Anichkina. Limit missile defense-or expand it? : a Russian response. Bulletin of the Atomic Scientists, 71(2):17–20, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:APC

[Ano45a] Anonymous. Atomic power control problems. Bulletin of the Atomic Scientists of Chicago, 1(2):1, 5, December 24, 1945. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:ASC

[Ano45b] Anonymous. The Atomic Scientists of Chicago. Bulletin of the Atomic Scientists of Chicago, 1(1):1, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:CAL

[Ano45c] Anonymous. Chicago activities …: A book on the A-bomb in preparation; Speaker’s Bureau; Radio; Materials Committee; Records Committee; Office of Enquiry of the University; I.C.C. Science Section. Bulletin of the Atomic Scientists of Chicago, 1 (1):6, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:CPR

[Ano45d] Anonymous. Conference with press representatives. Bulletin of the Atomic Scientists of Chicago, 1(2):5, December 24, 1945. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:CNS

[Ano45e] Anonymous. Congressional news: Senate hearings on atomic energy. Bulletin of the Atomic Scientists of Chicago, 1(2):3–4, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:DCA

[Ano45f] Anonymous. Denver conference on atomic energy. Bulletin of the Atomic Scientists of Chicago, 1(2):5, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:FASa

[Ano45g] Anonymous. The Federation of Atomic Scientists. Bulletin of the Atomic Scientists of Chicago, 1(1):2, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:FASb

[Ano45h] Anonymous. The Federation of Atomic Scientists: National Committee on Atomic Information; National Organization of Scientists. Bulletin of the Atomic Scientists of Chicago, 1(2):4, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:GGF

[Ano45i] Anonymous. General Groves on the future of the atomic bomb projects. Bulletin of the Atomic Scientists of Chicago, 1(2):2, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:CCS

[Ano45j] Anonymous. I. C. C. Science Section. Bulletin of the Atomic Scientists of Chicago, 1(2):5, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:MJB

[Ano45k] Anonymous. The May–Johnson Bill. Bulletin of the Atomic Scientists of Chicago, 1(1):3, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:MSH

[Ano45l] Anonymous. More on Senate hearings. Bulletin of the Atomic Scientists of Chicago, 1(2):5, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:NSS

[Ano45m] Anonymous. News of scientific societies: National Organization of Scientists; American Physical Society; French Academy of Sciences; American Chemical Society; National Academy of Sciences. Bulletin of the Atomic Scientists of Chicago, 1(1):5, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:PAA

[Ano45n] Anonymous. The policy of the A. L. A. S. [Association of Los Alamos Scientists]. Bulletin of the Atomic Scientists of Chicago, 1(1):2, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the A. L. A. S. Newsletter.

Anonymous:1945:SHA

[Ano45o] Anonymous. Senate hearings on atomic energy. Bulletin of the Atomic Scientists of Chicago, 1(1):3–4, December 10, 1945. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:TFA

[Ano45p] Anonymous. Technical feasability of atomic energy controls. Bulletin of the Atomic Scientists of Chicago, 1(2):1, December 24, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1945:TAK

[Ano45q] Anonymous. Truman–Atlee–King conference. Bulletin of the Atomic Scientists of Chicago, 1(1):4, December 10, 1945. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:TLB

[Ano46a] Anonymous. 70 thousand letters back the McMahon Bill. Bulletin of the Atomic Scientists, 1(9):6, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ARP

[Ano46b] Anonymous. The American and Russian proposals …. Bulletin of the Atomic Scientists, 2(1–2):1, 24, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ASR

[Ano46c] Anonymous. Applied science restricted in Germany. Bulletin of the Atomic Scientists, 1(10):8, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:APB

[Ano46d] Anonymous. ASC publishes booklet The Atomic Bomb. Bulletin of the Atomic Scientists of Chicago, 1(5):11, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ANC

[Ano46e] Anonymous. Association of Northern California Scientists. Bulletin of the Atomic Scientists, 1(8):18, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ABO

[Ano46f] Anonymous. The atomic bomb and our cities … from report of U.S. Strategic Bombing Survey. Bulletin of the Atomic Scientists, 2(3–4):29–30, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ABTd

[Ano46g] Anonymous. Atomic bomb tests: A land test? Bulletin of the Atomic Scientists, 1(9):18, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ABTb

[Ano46h] Anonymous. Atomic bomb tests: Bikini test plans. Bulletin of the Atomic Scientists, 1(9):18, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ABTa

[Ano46i] Anonymous. Atomic bomb tests: First power pile to be built at Clinton. Bulletin of the Atomic Scientists, 1(9):18, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ABTe

[Ano46j] Anonymous. Atomic bomb tests: Lie prods Atomic Energy Commission members. Bulletin of the Atomic Scientists, 1(9):18, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ABTc

[Ano46k] Anonymous. Atomic bomb tests: Navy announces bomb test evaluators. Bulletin of the Atomic Scientists, 1(9):18, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AEA

[Ano46l] Anonymous. The Atomic Energy Act of 1946. Bulletin of the Atomic Scientists, 2(3–4):18–25, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AECa

[Ano46m] Anonymous. Atomic energy controls no. 5 — control of mineral production. Bulletin of the Atomic Scientists of Chicago, 1(6): 4, 9, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ARI

[Ano46n] Anonymous. The availability of radioactive isotopes … announced by the Manhattan Project. Bulletin of the Atomic Scientists, 2(1–2):20–22, 24, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BHD

[Ano46o] Anonymous. Baruch and Hancock defend the American plan. Bulletin of the Atomic Scientists, 2(9–10):20–23, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BAA

[Ano46p] Anonymous. Baruch appointed to UNO Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(8):20, April 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BTR

[Ano46q] Anonymous. The Bikini tests — radiological effects. Bulletin of the Atomic Scientists, 2(3–4):26, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BPU

[Ano46r] Anonymous. Bill permits user of surplus sales for international exchange of students. Bulletin of the Atomic Scientists, 2(5– 6):16–, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BMH

[Ano46s] Anonymous. Bomb making holiday suggested. Bulletin of the Atomic Scientists, 1(9):20, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BPP

[Ano46t] Anonymous. “Look” publishes pictorial summary of atomic bomb. Bulletin of the Atomic Scientists of Chicago, 1(6):12, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BWN

[Ano46u] Anonymous. “One World or None” to be issued soon. Bulletin of the Atomic Scientists of Chicago, 1(6):12, March 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BAW

[Ano46v] Anonymous. Our Atomic World published. Bulletin of the Atomic Scientists, 2(5–6):24, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BAE

[Ano46w] Anonymous. The British Atomic Energy Project. Bulletin of the Atomic Scientists, 1(9):19, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BFP

[Ano46x] Anonymous. British foreign policy and the bomb. Bulletin of the Atomic Scientists, 1(7):15, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:BVA

[Ano46y] Anonymous. British views of atomic energy … from Nature. Bulletin of the Atomic Scientists, 2(3–4):13–15, August 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CAE

[Ano46z] Anonymous. The Canadian Atomic Energy Project. Bulletin of the Atomic Scientists, 1(7):7, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CAC

[Ano46-27] Anonymous. Chairman of American Chemical Society calls for international cooperation of scientists. Bulletin of the Atomic Scientists, 1(10):14, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CLT

[Ano46-28] Anonymous. Clinton Laboratories training program. Bulletin of the Atomic Scientists, 2(7–8):28, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CFC

[Ano46-29] Anonymous. Committee for Foreign Correspondence. Bulletin of the Atomic Scientists, 2(3–4):26, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CSB

[Ano46-30] Anonymous. A compromised suggestion by the British Association of Scientific Workers. Bulletin of the Atomic Scientists, 2 (7–8):29, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CNG

[Ano46-31] Anonymous. Congressional news: General Groves presents his views to Senate Committee. Bulletin of the Atomic Scientists of Chicago, 1(6):5, 8, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CNS

[Ano46-32] Anonymous. Congressional news: Senate hearings on atomic energy. Bulletin of the Atomic Scientists of Chicago, 1(5):5, 8–9, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CAS

[Ano46-33] Anonymous. The Cornell Association of Scientists. Bulletin of the Atomic Scientists, 1(10):16, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:C

[Ano46-34] Anonymous. Correction. Bulletin of the Atomic Scientists, 2(3– 4):25, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CFA

[Ano46-35] Anonymous. Council of Federation of American Scientists meets. Bulletin of the Atomic Scientists, 1(10):16, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CSD

[Ano46-36] Anonymous. Current status of domestic legislation. Bulletin of the Atomic Scientists, 1(8):1, 19, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:DL

[Ano46-37] Anonymous. A dangerous lull. Bulletin of the Atomic Scientists, 1 (11):1, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:DCI

[Ano46-38] Anonymous. Declassification committee issues report. Bulletin of the Atomic Scientists of Chicago, 1(5):11, February 15, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:EAP

[Ano46-39] Anonymous. Errata: “Atomic Power Production”, by Farrington Daniels. Bulletin of the Atomic Scientists, 1(8):18, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:EUN

[Ano46-40] Anonymous. Errata: “Uranium in Nature”. Bulletin of the Atomic Scientists, 1(8):18, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:FCM

[Ano46-41] Anonymous. FAmS Council meets. Bulletin of the Atomic Scientists, 1(9):20, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:FAS

[Ano46-42] Anonymous. Federation of Atomic Scientists. Bulletin of the Atomic Scientists, 2(5–6):24, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:FMA

[Ano46-43] Anonymous. First member of the UNO Atomic Commission appointed. Bulletin of the Atomic Scientists, 1(7):15, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:HBS

[Ano46-44] Anonymous. Hanson Baldwin on secrecy provisions. Bulletin of the Atomic Scientists, 1(9):6, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:CCR

[Ano46-45] Anonymous. I.C.C. rally backs civilian control of atomic energy. Bulletin of the Atomic Scientists of Chicago, 1(6):11, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ICU

[Ano46-46] Anonymous. Intercollegiate conference urges UN atomic rule. Bulletin of the Atomic Scientists, 1(9):12, April 15, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:IAS

[Ano46-47] Anonymous. Iowa Academy of Sciences supports Lilienthal– Acheson plan. Bulletin of the Atomic Scientists, 1(12):15, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:KM

[Ano46-48] Anonymous. Kansas meetings. Bulletin of the Atomic Scientists, 1(9):20, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:LLA

[Ano46-49] Anonymous. Labor leaders and atomic scientists meet. Bulletin of the Atomic Scientists, 1(8):18, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:LWV

[Ano46-50] Anonymous. League of Women Voters puts atomic energy first. Bulletin of the Atomic Scientists, 1(10):11, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:LMPa

[Ano46-51] Anonymous. A list of Manhattan Project declassified documents: Part A. Bulletin of the Atomic Scientists, 2(11–12):15–16, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:LMPb

[Ano46-52] Anonymous. A list of Manhattan Project declassified documents: Part B. Bulletin of the Atomic Scientists, 2(11–12):16–17, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:LCS

[Ano46-53] Anonymous. London conference [: “Science and welfare of mankind”]. Bulletin of the Atomic Scientists, 1(7):7, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MBSa

[Ano46-54] Anonymous. The McMahon Bill and the State Department report. Bulletin of the Atomic Scientists, 1(9):1, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MBP

[Ano46-55] Anonymous. McMahon Bill passes the Senate. Bulletin of the Atomic Scientists, 1(12):11, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MBSb

[Ano46-56] Anonymous. McMahon Bill status discussed by Chemical and Engineering News. Bulletin of the Atomic Scientists, 1(9):6, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MCH

[Ano46-57] Anonymous. The McMahon Committee hearings on international control delayed. Bulletin of the Atomic Scientists, 1(12):11, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MIU

[Ano46-58] Anonymous. Medical and industrial uses of pile products. Bulletin of the Atomic Scientists, 1(9):1, 16–17, April 15, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from The Atomic Bomb, by the Atomic Scientists of Chicago, Chicago, 1946.

Anonymous:1946:MCR

[Ano46-59] Anonymous. Midwest conference with religious leaders. Bulletin of the Atomic Scientists of Chicago, 1(5):12, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MCC

[Ano46-60] Anonymous. Military or civilian control of atomic energy? Bulletin of the Atomic Scientists, 1(9):1, 16, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:MDS

[Ano46-61] Anonymous. Momentous decision: Senate Committee prepares domestic bill. Bulletin of the Atomic Scientists of Chicago, 1(6): 1, 12, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:NCU

[Ano46-62] Anonymous. National Commission of UNESCO named. Bulletin of the Atomic Scientists, 2(9–10):25, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:NCC

[Ano46-63] Anonymous. National Committee for Civilian Control announced. Bulletin of the Atomic Scientists, 1(8):18, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:NTB

[Ano46-64] Anonymous. Navy test [on Bikini Atoll] postponed. Bulletin of the Atomic Scientists, 1(8):20, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:NEC

[Ano46-65] Anonymous. New Executive Committee of the Atomic Scientists of Chicago. Bulletin of the Atomic Scientists, 1(12):15, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:NGA

[Ano46-66] Anonymous. New government agencies … Atomic Energy Commission appointed. Bulletin of the Atomic Scientists, 2(9–10):24, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). AEC members appointed by President Truman are David E. Lilienthal (Chairman), Robert F. Bacher, Sumner T. Pike, Lewis L. Strauss, and William W. Waymack.

Anonymous:1946:NIA

[Ano46-67] Anonymous. No insurance for atomic bomb victims. Bulletin of the Atomic Scientists, 1(8):18, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:NRC

[Ano46-68] Anonymous. Nuclear research center at Camp Upton, N.Y. Bulletin of the Atomic Scientists, 2(5–6):21, September 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ORI

[Ano46-69] Anonymous. The Oak Ridge Institute of Nuclear Studies. Bulletin of the Atomic Scientists, 2(5–6):21, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:OLB

[Ano46-70] Anonymous. Obituary: Louis B. Slotin. Bulletin of the Atomic Scientists, 1(12):16, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://en.

wikipedia.org/wiki/Criticality_accident.

Anonymous:1946:OCE

[Ano46-71] Anonymous. “Operation Crossroads”: The effect of the atomic bomb on naval power. Bulletin of the Atomic Scientists of Chicago, 1(5):1, 12, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PIC

[Ano46-72] Anonymous. Plans for international control take form. Bulletin of the Atomic Scientists, 1(8):1, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PNR

[Ano46-73] Anonymous. Plans for nuclear research in U.S. … the Argonne National Laboratory. Bulletin of the Atomic Scientists, 2(5–6):20, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PCR

[Ano46-74] Anonymous. Power costs reported to UN Commission. Bulletin of the Atomic Scientists, 2(5–6):17, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PSD

[Ano46-75] Anonymous. The present status of declassification. Bulletin of the Atomic Scientists, 2(11–12):14, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PES

[Ano46-76] Anonymous. President establishes Scientific Research Board. Bulletin of the Atomic Scientists, 2(9–10):25, November 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PTS

[Ano46-77] Anonymous. President Truman’s speech at Fordham. Bulletin of the Atomic Scientists, 1(11):6, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PAU

[Ano46-78] Anonymous. Prof. Auger on U.S. control plan. Bulletin of the Atomic Scientists, 2(5–6):17, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:PBO

[Ano46-79] Anonymous. Publications of the Baruch Office. Bulletin of the Atomic Scientists, 2(7–8):32, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:QWE

[Ano46-80] Anonymous. Quincy Wright and Edith Wynner address the ASC. Bulletin of the Atomic Scientists of Chicago, 1(5):12, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:RSD

[Ano46-81] Anonymous. Repercussions of the State Dep’t report: Opinions of the Lilienthal board members. Bulletin of the Atomic Scientists, 1 (9):11, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:RSC

[Ano46-82] Anonymous. Reports of Senate Committee to be republished abroad. Bulletin of the Atomic Scientists, 1(9):6, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:RPF

[Ano46-83] Anonymous. Resolutions passed by Federation of American Scientists. Bulletin of the Atomic Scientists, 2(1–2):23, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:RRC

[Ano46-84] Anonymous. Resolutions request cessation of bomb production. Bulletin of the Atomic Scientists, 1(7):14, March 15, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:RMB

[Ano46-85] Anonymous. The revised McMahon Bill. Bulletin of the Atomic Scientists, 1(9):2–5, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:RCC

[Ano46-86] Anonymous. Rollins College conference. Bulletin of the Atomic Scientists, 1(7):14, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Conference “to discuss and recommend the formation of a type of world government adequate for the control of atomic bombs and other super-weapons.”.

Anonymous:1946:RAB

[Ano46-87] Anonymous. Russia and the atomic bomb. Bulletin of the Atomic Scientists of Chicago, 1(5):10–11, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SCSb

[Ano46-88] Anonymous. Scientists comment on State Dep’t report: American Chemical Society recommends the report. Bulletin of the Atomic Scientists, 1(9):12, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SCSa

[Ano46-89] Anonymous. Scientists comment on State Dep’t report: Statement by the Federation of American Scientists. Bulletin of the Atomic Scientists, 1(9):12, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SRM

[Ano46-90] Anonymous. Second Rocky Mountain Conference [on Atomic Energy, Estes Park, Colorado, 24–26 June 1946]. Bulletin of the Atomic Scientists, 2(3–4):25, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SWP

[Ano46-91] Anonymous. Secretary of War Patterson favors the McMahon Bill — with amendments. Bulletin of the Atomic Scientists of Chicago, 1(6):6–8, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SHA

[Ano46-92] Anonymous. Senate hearings on atomic energy. Bulletin of the Atomic Scientists, 1(7):10, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SBC

[Ano46-93] Anonymous. Senator Ball on control of atomic energy. Bulletin of the Atomic Scientists, 1(8):20, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SRN

[Ano46-94] Anonymous. Staff reorganization in the National Committee for Atomic Information. Bulletin of the Atomic Scientists, 2(5–6):23, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SDR

[Ano46-95] Anonymous. State Dep’t report on the control of atomic energy. Bulletin of the Atomic Scientists, 1(8):2–9, April 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:SDL

[Ano46-96] Anonymous. The status of domestic legislation …. Bulletin of the Atomic Scientists, 2(1–2):23, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:S

[Ano46-97] Anonymous. Structure of the UNO. Bulletin of the Atomic Scientists of Chicago, 1(3):1, 4, January 10, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AB

[Ano46-98] Anonymous. “The atomic bomb”. Bulletin of the Atomic Scientists, 1(8):20, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Second edition of ASC pamphlet.

Anonymous:1946:TAC

[Ano46-99] Anonymous. Truman appoints committee to report on bomb tests. Bulletin of the Atomic Scientists, 1(8):20, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:ACA

[Ano46-100] Anonymous. UN Atomic Commission almost complete. Bulletin of the Atomic Scientists, 1(10):8, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AECc

[Ano46-101] Anonymous. The UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 2(3–4):2–12, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AECd

[Ano46-102] Anonymous. The UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 2(5–6):12–13, 16, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AECe

[Ano46-103] Anonymous. The UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 2(11–12):13, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:AECb

[Ano46-104] Anonymous. UN Atomic Energy Commission called together. Bulletin of the Atomic Scientists, 1(12):11, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:UNA

[Ano46-105] Anonymous. The United Nations and atomic energy. Bulletin of the Atomic Scientists, 2(9–10):16–19, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:FPA

[Ano46-106] Anonymous. UNO faces the problem of atomic energy control. Bulletin of the Atomic Scientists of Chicago, 1(3):1, 3, January 10, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:UAB

[Ano46-107] Anonymous. Use of atomic bomb assailed by [Msrg Fulton J.] Sheen. Bulletin of the Atomic Scientists, 1(10):12, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1946:WFS

[Ano46-108] Anonymous. A world federation of scientific workers. Bulletin of the Atomic Scientists, 2(5–6):7, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:ARS

[Ano47a] Anonymous. AAAS resolution on science legislation. Bulletin of the Atomic Scientists, 3(2):36, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:AENb

[Ano47b] Anonymous. Atomic energy in the news. Bulletin of the Atomic Scientists, 3(7):199–200, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:AENa

[Ano47c] Anonymous. Atomic energy in the news … in the U.S. Senate; in the United Nations; in the U.S. Atomic Energy Commission. Bulletin of the Atomic Scientists, 3(4–5):121, April/May 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:BAS

[Ano47d] Anonymous. British atomic scientists’ proposals for international control of atomic energy. Bulletin of the Atomic Scientists, 3(2): 42–43, 49, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:CBA

[Ano47e] Anonymous. Comparison of British and American Atomic Energy Acts. Bulletin of the Atomic Scientists, 3(2):50–51, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:DAE

[Ano47f] Anonymous. Disarmament and atomic energy in the Security Council: Summary of the Security Council’s deliberations on atomic energy and of Gromyko’s March 5 speech. Bulletin of the Atomic Scientists, 3(3):86–91, 97, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:DDG

[Ano47g] Anonymous. Disarmament debate in the UN General Assembly. Bulletin of the Atomic Scientists, 3(1):6–8, January 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:FFS

[Ano47h] Anonymous. FAS forms a Scientists’ Committee on Loyalty. Bulletin of the Atomic Scientists, 3(12):342, 347, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:FUI

[Ano47i] Anonymous. FAS urges international office for isotope distribution. Bulletin of the Atomic Scientists, 3(12):248, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:FRA

[Ano47j] Anonymous. First report of the Atomic Energy Commission to the Security Council. Bulletin of the Atomic Scientists, 3(1):16– 27, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:ICA

[Ano47k] Anonymous. International control of atomic energy: selections from the Bulletin of the Atomic Scientists. Atomic Scientists of Chicago, Chicago, IL, USA, 1947. 46 pp. LCCN ????

Anonymous:1947:JHI

[Ano47l] Anonymous. Just how important is the atomic bomb? Bulletin of the Atomic Scientists, 3(10):313, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:LMP

[Ano47m] Anonymous. List of Manhattan Project declassified documents — II. Bulletin of the Atomic Scientists, 3(1):28–30, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:MPS

[Ano47n] Anonymous. McCloy predicts super atomic bombs within decade. Bulletin of the Atomic Scientists, 3(1):5, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:NSF

[Ano47o] Anonymous. The National Science Foundation Bill and the President’s veto. Bulletin of the Atomic Scientists, 3(10):290, 300, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:PCS

[Ano47p] Anonymous. President Conant supports National Science Foundation. Bulletin of the Atomic Scientists, 3(1):2, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:PGG

[Ano47q] Anonymous. Principles governing the general regulation and reduction of armaments. Bulletin of the Atomic Scientists, 3(1): 9, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:PAE

[Ano47r] Anonymous. Proceedings of the UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 3(1):10–15, January 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:SRC

[Ano47s] Anonymous. The second report of the UN A.E.C. to the Security Council. Bulletin of the Atomic Scientists, 3(11):332–334, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:SCAa

[Ano47t] Anonymous. Senate Committee on Atomic Energy. Bulletin of the Atomic Scientists, 3(2):44–47, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:SCAb

[Ano47u] Anonymous. Senate Committee on Atomic Energy interrogates Lilienthal and other commissioners. Bulletin of the Atomic Scientists, 3(3):76–78, 94–96, 98, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:SDM

[Ano47v] Anonymous. The Senate debates Mr. Lilienthal’s confirmation: March 24–April 9, 1947. Bulletin of the Atomic Scientists, 3(6): 158–162, 168, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:SSA

[Ano47w] Anonymous. Soviet statements on atomic energy and U.S. answers: Harold Stassen interviews Joseph Stalin: American answers by F. H. Osborn and F. A. Lindsay and Senator McMahon. Bulletin of the Atomic Scientists, 3(7):187–192, July 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:DBL

[Ano47x] Anonymous. That I deeply believe: Lilienthal defines democracy for Senator McKellar. Bulletin of the Atomic Scientists, 3(3):78, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:UNA

[Ano47y] Anonymous. United Nations and atomic energy. Bulletin of the Atomic Scientists, 3(2):55–56, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). A report from London of the Parliamentary debate on the British Atomic Energy Act of 1946, and subsequent press comment.

Anonymous:1947:USA

[Ano47z] Anonymous. United States Atomic Energy Commission reports to Congress. Bulletin of the Atomic Scientists, 3(9):275–280, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:UOA

[Ano47-27] Anonymous. U.S. organization for atomic energy control: Personnel of U.S. Atomic Energy Organization. Bulletin of the Atomic Scientists, 3(3):79, 82, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:WDT

[Ano47-28] Anonymous. War Department thinking on the atomic bomb. Bulletin of the Atomic Scientists, 3(6):150–155, 168, June 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1947:WDR

[Ano47-29] Anonymous. World distribution of radioisotopes announced. Bulletin of the Atomic Scientists, 3(10):309, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AIP

[Ano48a] Anonymous. AEC interim procedure for local security boards. Bulletin of the Atomic Scientists, 4(7):198, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:APP

[Ano48b] Anonymous. AEC proposed patent compensation regulations. Bulletin of the Atomic Scientists, 4(8):244, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:CTE

[Ano48c] Anonymous. A.E.C. terms extended for two years. Bulletin of the Atomic Scientists, 4(8):248, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AECb

[Ano48d] Anonymous. AEC to establish chemical laboratory [in New Brunswick, NJ]. Bulletin of the Atomic Scientists, 4(8):252, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:ASN

[Ano48e] Anonymous. American Society of Newspaper Editors reports on atomic information problems. Bulletin of the Atomic Scientists, 4(7):211–212, 217, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:ASA

[Ano48f] Anonymous. Association of Scientists for Atomic Education News. Bulletin of the Atomic Scientists, 4(1):32, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AEB

[Ano48g] Anonymous. Atomic energy — 1948: A Business Week report. Bulletin of the Atomic Scientists, 4(7):213–217, July 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from 10 April 1948 issue of Business Week.

Anonymous:1948:AENd

[Ano48h] Anonymous. Atomic energy and the national party platforms. Bulletin of the Atomic Scientists, 4(8):234, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AECa

[Ano48i] Anonymous. The Atomic Energy Commission reports to the Congress. Bulletin of the Atomic Scientists, 4(3):88–96, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AED

[Ano48j] Anonymous. Atomic energy debates in the UN. Bulletin of the Atomic Scientists, 4(11):326–334, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AEG

[Ano48k] Anonymous. Atomic energy in Goebbels’ diary. Bulletin of the Atomic Scientists, 4(5):134, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AENa

[Ano48l] Anonymous. Atomic energy in the news. Bulletin of the Atomic Scientists, 4(3):76, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AENb

[Ano48m] Anonymous. Atomic energy in the news. Bulletin of the Atomic Scientists, 4(4):110, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AENc

[Ano48n] Anonymous. Atomic energy in the news: Senator Wherry moves to restore military control; search for radioactive ores in Australia; notes on the defense of Dr. Condon; announcements from the AEC. Bulletin of the Atomic Scientists, 4(5):138, May 1948.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:AEA

[Ano48o] Anonymous. Attack is extended to atomic physics. Bulletin of the Atomic Scientists, 4(12):371, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:BNF

[Ano48p] Anonymous. Bibliography on nuclear fission as a source of power. Bulletin of the Atomic Scientists, 4(4):121, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:BEA

[Ano48q] Anonymous. A bill to extend AEC terms for two years. Bulletin of the Atomic Scientists, 4(6):162, 180, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:BRC

[Ano48r] Anonymous. Bulletin requests clarification from Governor Dewey. Bulletin of the Atomic Scientists, 4(10):298, October 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Wire to Governor Dewey’s train on 27 September 1948.

Anonymous:1948:BSR

[Ano48s] Anonymous. The N. Y. Sun retracts inaccurate statements about the Bulletin. Bulletin of the Atomic Scientists, 4(2):63, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:CPO

[Ano48t] Anonymous. The charges presented in Oak Ridge cases. Bulletin of the Atomic Scientists, 4(7):196, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:CAA

[Ano48u] Anonymous. The Committee on Un-American Activities feels the effect of November 2 [Congressional elections of 1948]. Bulletin of the Atomic Scientists, 4(12):376, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:CFS

[Ano48v] Anonymous. Communications from French scientists. Bulletin of the Atomic Scientists, 4(10):309–311, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a, Ash48, Rit48, Ler48].

Anonymous:1948:CCA

[Ano48w] Anonymous. Condon is cleared by Atomic Energy Commission. Bulletin of the Atomic Scientists, 4(8):226, 255, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:DEL

[Ano48x] Anonymous. Dr. Einstein’s letter. Bulletin of the Atomic Scientists, 4(11):332, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ein48a, Ano48-33].

Anonymous:1948:EBR

[Ano48y] Anonymous. Erratum: Book review. Bulletin of the Atomic Scientists, 4(3):79, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:FUI

[Ano48z] Anonymous. FAS urges international office for isotope distribution. Bulletin of the Atomic Scientists, 4(8):248, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:FAP

[Ano48-27] Anonymous. Fifth anniversary of the pile. Bulletin of the Atomic Scientists, 4(1):32, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The plaque at Stagg Field at the University of Chicago reads: “On December 2, 1942 Man Achieved Here The First Self-Sustaining Chain Reaction And Thereby Initiated the Controlled Release of Nuclear Energy”.

Anonymous:1948:FSAa

[Ano48-28] Anonymous. Fourth semi-annual report of the United States Atomic Energy Commission. [part 1]. Bulletin of the Atomic Scientists, 4(9):268–270, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Part 2 [Ano48-29].

Anonymous:1948:FSAb

[Ano48-29] Anonymous. Fourth semi-annual report of the United States Atomic Energy Commission. [part 2]. Bulletin of the Atomic Scientists, 4(11):348–352, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Part 1 [Ano48-28].

Anonymous:1948:GDA

[Ano48-30] Anonymous. Governor Dewey and the atom. Bulletin of the Atomic Scientists, 4(10):297–298, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint of New York Times editorial of 25 September 1948.

Anonymous:1948:IMA

[Ano48-31] Anonymous. An international moratorium on atomic energy for power uses. Bulletin of the Atomic Scientists, 4(6):183–184, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:NSC

[Ano48-32] Anonymous. New strategic concept of defense of U.S. Bulletin of the Atomic Scientists, 4(4):127–128, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:PVE

[Ano48-33] Anonymous. Polish version of Einstein’s message. Bulletin of the Atomic Scientists, 4(10):320, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The article begins: “This is a translation of the message attributed to Dr. Einstein in Poland. It is completely dissimilar to the actual message [Ein48a]. See explanation [Ano48x].

Anonymous:1948:PTU

[Ano48-34] Anonymous. President Truman upholds the Atomic Energy Act. Bulletin of the Atomic Scientists, 4(9):267, 288, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:PCA

[Ano48-35] Anonymous. Princeton Conference of Atomic Scientists. Bulletin of the Atomic Scientists, 4(1):32, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:RCJ

[Ano48-36] Anonymous. Report of the Congressional Joint Committee. Bulletin of the Atomic Scientists, 4(3):96, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:SVR

[Ano48-37] Anonymous. Senator Vandenberg reaffirms support of civilian control. Bulletin of the Atomic Scientists, 4(10):299, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:SIR

[Ano48-38] Anonymous. Stable isotopes released by A. E. C. Bulletin of the Atomic Scientists, 4(1):13, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:TRU

[Ano48-39] Anonymous. Third report of the United Nations Atomic Energy Commission. Bulletin of the Atomic Scientists, 4(6):181–182, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:CRS

[Ano48-40] Anonymous. UN Commission rejects Soviet control proposals. Bulletin of the Atomic Scientists, 4(5):129–130, 156–158, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1948:UAE

[Ano48-41] Anonymous. U.S. Atomic Energy Commission. Bulletin of the Atomic Scientists, 4(1):14, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:AAC

[Ano49a] Anonymous. AEC announces changes in personnel. Bulletin of the Atomic Scientists, 5(6–7):203, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:ACS

[Ano49b] Anonymous. AEC criteria for security clearance. Bulletin of the Atomic Scientists, 5(2):62–63, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:AFSd

[Ano49c] Anonymous. AEC fifth semiannual report: Part II. Bulletin of the Atomic Scientists, 5(4):114–125, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:AFSc

[Ano49d] Anonymous. AEC fifth semiannual report: Sections I and II. Bulletin of the Atomic Scientists, 5(3):83–93, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:ASC

[Ano49e] Anonymous. AEC signs contract with Western Electric for operation of Sandia. Bulletin of the Atomic Scientists, 5(11):324, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:AIR

[Ano49f] Anonymous. AEC to issue report on weapon effects. Bulletin of the Atomic Scientists, 5(3):68, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:ABN

[Ano49g] Anonymous. Announcement: [birth of the nucleonic instruments industry]. Bulletin of the Atomic Scientists, 5(12):336, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:BHD

[Ano49h] Anonymous. Booklet on how to do business with AEC now available. Bulletin of the Atomic Scientists, 5(11):308, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:BR

[Ano49i] Anonymous. Books received. Bulletin of the Atomic Scientists, 5(12):354, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:BASl

[Ano49j] Anonymous. Bulletin of the Atomic Scientists: Index to volume 5. Bulletin of the Atomic Scientists, 5(12):355–356, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:BASb

[Ano49k] Anonymous. British Atomic Scientists annual conference. Bulletin of the Atomic Scientists, 5(2):57–62, February 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpted from Atomic Scientists News, number 11, 7 January 1949.

Anonymous:1949:CLS

[Ano49l] Anonymous. Civil liberties of scientists: Report of AAAS Committee. Bulletin of the Atomic Scientists, 5(11):298–299, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:CWR

[Ano49m] Anonymous. Commissioner Waymack resigns. Bulletin of the Atomic Scientists, 5(1):32, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:DSC

[Ano49n] Anonymous. DuPont to survey chemical processes in plutonium production. Bulletin of the Atomic Scientists, 5(2):64, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:E

[Ano49o] Anonymous. Erratum. Bulletin of the Atomic Scientists, 5(4): 128, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:FAPf

[Ano49p] Anonymous. FAS announces policy decisions. Bulletin of the Atomic Scientists, 5(6–7):184–185, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:FSW

[Ano49q] Anonymous. FAS suggests working control plan to UNAEC. Bulletin of the Atomic Scientists, 5(3):96, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:FAPa

[Ano49r] Anonymous. French atomic pile. Bulletin of the Atomic Scientists, 5(1):32, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:GIL

[Ano49s] Anonymous. The great inquiry lapses. Bulletin of the Atomic Scientists, 5(10):282, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:GIT

[Ano49t] Anonymous. The great inquiry: Testimony at AEC hearings. Bulletin of the Atomic Scientists, 5(8–9):221–250, 254, August/ September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This report is a condensation of the 2000page transcript.

Anonymous:1949:HND

[Ano49u] Anonymous. Hafstad named Director of Reactor Development Division. Bulletin of the Atomic Scientists, 5(2):47, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:HGC

[Ano49v] Anonymous. History of the genetics conflict. Bulletin of the Atomic Scientists, 5(5):131–140, 156, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:HHG

[Ano49w] Anonymous. Hyman H. Goldsmith. Bulletin of the Atomic Scientists, 5(8–9):206, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the introduction: “Hyman H. Goldsmith, the initiator of the Bulletin of the Atomic Scientists and its co-editor, died on August 7, 1949, at South Wyndham, Vermont, as the result of an accident.”.

Anonymous:1949:IRS

[Ano49x] Anonymous. Interim report on six-power talks. Bulletin of the Atomic Scientists, 5(12):347–351, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:ID

[Ano49y] Anonymous. Isotope distribution. Bulletin of the Atomic Scientists, 5(12):350, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:JCR

[Ano49z] Anonymous. Joint Committee’s reports on AEC investigation. Bulletin of the Atomic Scientists, 5(12):330–336, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:JCA

[Ano49-27] Anonymous. Joliot-Curie and the atomic secrets. Bulletin of the Atomic Scientists, 5(4):109–110, 126, April 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from The Economist.

Anonymous:1949:LCC

[Ano49-28] Anonymous. Lilienthal and Commission cleared. Bulletin of the Atomic Scientists, 5(11):300, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:MEP

[Ano49-29] Anonymous. Major expansion planned for U.S. atomic plants. Bulletin of the Atomic Scientists, 5(11):324, November 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:MPC

[Ano49-30] Anonymous. Military and political consequences of atomic energy: Some views on Blackett’s book. Bulletin of the Atomic Scientists, 5(2):34, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:MLR

[Ano49-31] Anonymous. Mr. Lilienthal resigns. Bulletin of the Atomic Scientists, 5(12):328, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:PND

[Ano49-32] Anonymous. Pitzer new director AEC Research Division. Bulletin of the Atomic Scientists, 5(1):32, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:RTG

[Ano49-33] Anonymous. Radioisotope training given to foreign scientists at Oak Ridge. Bulletin of the Atomic Scientists, 5(2):64, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:REM

[Ano49-34] Anonymous. The Russian explosion: Mr. Truman’s announcement. Bulletin of the Atomic Scientists, 5(10):261, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:SOW

[Ano49-35] Anonymous. Scientists objecting to war work organize. Bulletin of the Atomic Scientists, 5(11):312, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:SDJ

[Ano49-36] Anonymous. Secrecy debated in Joint Congressional Committee. Bulletin of the Atomic Scientists, 5(3):95, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:SPA

[Ano49-37] Anonymous. Security problems for the administrators’ viewpoint. Bulletin of the Atomic Scientists, 5(12):351, December 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:SA

[Ano49-38] Anonymous. The Soviet announcement. Bulletin of the Atomic Scientists, 5(10):261, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:UAE

[Ano49-39] Anonymous. U.S. Atomic Energy Commission organization chart. Bulletin of the Atomic Scientists, 5(4):112–113, April 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:USB

[Ano49-40] Anonymous. U.S. strategic bombing survey appraises use of atomic bomb. Bulletin of the Atomic Scientists, 5(11):324, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:WCR

[Ano49-41] Anonymous. Westinghouse to construct reactor for ship propulsion. Bulletin of the Atomic Scientists, 5(2):47, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1949:WCI

[Ano49-42] Anonymous. WFUNA calls for immediate disarmament. Bulletin of the Atomic Scientists, 5(1):16, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:AFI

[Ano50a] Anonymous. AEC fellowships in industrial medicine. Bulletin of the Atomic Scientists, 6(10):298, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:APF

[Ano50b] Anonymous. AEC Predoctoral Fellowship Program. Bulletin of the Atomic Scientists, 6(4):127, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:AAB

[Ano50c] Anonymous. The Atomic Age and the belief in progress. Bulletin of the Atomic Scientists, 6(11):331, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:BRa

[Ano50d] Anonymous. Books received. Bulletin of the Atomic Scientists, 6(5):160, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:BRb

[Ano50e] Anonymous. Books received. Bulletin of the Atomic Scientists, 6(7):222, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:BRc

[Ano50f] Anonymous. Books received. Bulletin of the Atomic Scientists, 6(10):317, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:BAS

[Ano50g] Anonymous. Bulletin of the Atomic Scientists: Index to volume 6, 1950. Bulletin of the Atomic Scientists, 6(12):383–384, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:CLE

[Ano50h] Anonymous. Chronological list of events in civil defense. Bulletin of the Atomic Scientists, 6(8–9):270, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Anonymous:1950:CDA

[Ano50i] Anonymous. Civil defense across the nation. Bulletin of the Atomic Scientists, 6(8–9):276–278, 288, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Anonymous:1950:CDNa

[Ano50j] Anonymous. Civil defense news. Bulletin of the Atomic Scientists, 6(10):315–317, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:CDNb

[Ano50k] Anonymous. Civil defense news. Bulletin of the Atomic Scientists, 6(11):338–340, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:CBH

[Ano50l] Anonymous. Conference on biological hazards of atomic energy. Bulletin of the Atomic Scientists, 6(10):298, October 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:CHB

[Ano50m] Anonymous. Correction: [The H-bomb and after]. Bulletin of the Atomic Scientists, 6(4):127, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor50b].

Anonymous:1950:CSA

[Ano50n] Anonymous. Correction to Szilard article. Bulletin of the Atomic Scientists, 6(5):152, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:CAF

[Ano50o] Anonymous. The curtailment of the AEC Fellowship Program. Bulletin of the Atomic Scientists, 6(2):34, 62–63, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:DD

[Ano50p] Anonymous. The dangers of declassification, 1651. Bulletin of the Atomic Scientists, 6(4):105, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:DUG

[Ano50q] Anonymous. Debates in U.N. General Assembly: Soviet Union reverses position on quotas. Bulletin of the Atomic Scientists, 6 (2):43–49, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:DKF

[Ano50r] Anonymous. Dr. Klaus Fuchs to stand trial for espionage. Bulletin of the Atomic Scientists, 6(3):68, 94, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:ESW

[Ano50s] Anonymous. Editorial: Secrets will out. Bulletin of the Atomic Scientists, 6(3):67–68, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:FPC

[Ano50t] Anonymous. FAS protest Commission’s ban on discussion. Bulletin of the Atomic Scientists, 6(5):132, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:HRI

[Ano50u] Anonymous. The Hafstad report: Interdepartmental Committee on Scientific Research and Development. Bulletin of the Atomic Scientists, 6(12):379–380, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:HUA

[Ano50v] Anonymous. Highlights of 1950 in U.S. atomic energy programs. Bulletin of the Atomic Scientists, 6(10):307–313, October 1950.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:LPH

[Ano50w] Anonymous. Loyalty provisions in H.R. 4846. Bulletin of the Atomic Scientists, 6(4):98, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:LTC

[Ano50x] Anonymous. Loyalty tests cause cut in AEC Fellowship Program. Bulletin of the Atomic Scientists, 6(1):32, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:LTS

[Ano50y] Anonymous. Loyalty tests for science students? Bulletin of the Atomic Scientists, 6(4):98, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:MSM

[Ano50z] Anonymous. Mobilization of scientific manpower. Bulletin of the Atomic Scientists, 6(10):302, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pas51].

Anonymous:1950:NSF

[Ano50-27] Anonymous. The National Science Foundation Act of 1950. Bulletin of the Atomic Scientists, 6(6):186–190, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NRB

[Ano50-28] Anonymous. New reactor at Brookhaven. Bulletin of the Atomic Scientists, 6(10):314, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NWR

[Ano50-29] Anonymous. New weapons: Report of the Secretary of Defense. Bulletin of the Atomic Scientists, 6(6):174–177, June 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NBb

[Ano50-30] Anonymous. News in brief. Bulletin of the Atomic Scientists, 6 (4):128, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NBc

[Ano50-31] Anonymous. News in brief. Bulletin of the Atomic Scientists, 6(6):191–192, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NBd

[Ano50-32] Anonymous. News in brief. Bulletin of the Atomic Scientists, 6(7):220–222, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NBe

[Ano50-33] Anonymous. News in brief. Bulletin of the Atomic Scientists, 6(10):319, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NBf

[Ano50-34] Anonymous. News in brief. Bulletin of the Atomic Scientists, 6(11):351–352, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:NBA

[Ano50-35] Anonymous. News in brief: Activities of the U.S. Atomic Energy Commission; news from abroad. Bulletin of the Atomic Scientists, 6(2):64, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:OSM

[Ano50-36] Anonymous. The organization of scientific manpower: Reports of government advisory groups: The Barton report, American Institute of Physics. Bulletin of the Atomic Scientists, 6(12):355, 375–376, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:POE

[Ano50-37] Anonymous. The President orders exploration of the super bomb. Bulletin of the Atomic Scientists, 6(3):66, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:PTU

[Ano50-38] Anonymous. President Truman urges Science Foundation. Bulletin of the Atomic Scientists, 6(2):36, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:PTP

[Ano50-39] Anonymous. President Truman’s proposal for a new U.N. Disarmament Commission. Bulletin of the Atomic Scientists, 6(11): 343, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:PAR

[Ano50-40] Anonymous. Progress in AEC reactor program. Bulletin of the Atomic Scientists, 6(1):31–32, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:RNM

[Ano50-41] Anonymous. Review of new manuals on civilian protection: Atomic Attack: A Manuals for Survival, by John L. Balderson, Jr. and Gordon W. Hewes. Bulletin of the Atomic Scientists, 6 (11):340–341, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:RUP

[Ano50-42] Anonymous. Russian uranium procurement. Bulletin of the Atomic Scientists, 6(10):318, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:SFB

[Ano50-43] Anonymous. Science Foundation Bill passed by Congress. Bulletin of the Atomic Scientists, 6(5):132, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:SBL

[Ano50-44] Anonymous. Selected bibliography on the literature on civil defense. Bulletin of the Atomic Scientists, 6(8–9):272–275, August/ September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Anonymous:1950:STA

[Ano50-45] Anonymous. Senator Tydings asks disarmament conference. Bulletin of the Atomic Scientists, 6(3):79, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:SBD

[Ano50-46] Anonymous. “Super bomb” discussed. Bulletin of the Atomic Scientists, 6(2):36, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:TAD

[Ano50-47] Anonymous. Text of AEC directives. Bulletin of the Atomic Scientists, 6(5):132, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:TRS

[Ano50-48] Anonymous. The Trytten report: Scientific advisory committees to Selective Service Director. Bulletin of the Atomic Scientists, 6 (12):376–378, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1950:UBC

[Ano50-49] Anonymous. U.S.–British cooperation. Bulletin of the Atomic Scientists, 6(2):36, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:AMP

[Ano51a] Anonymous. AAAS manpower proposal. Bulletin of the Atomic Scientists, 7(4):118, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:ASS

[Ano51b] Anonymous. Atomic spies: A summary of the trials. Bulletin of the Atomic Scientists, 7(4):125–126, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:BR

[Ano51c] Anonymous. Books received. Bulletin of the Atomic Scientists, 7 (4):108, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:IVV

[Ano51d] Anonymous. Bulletin of the Atomic Scientists: Index to volume VII, 1951. Bulletin of the Atomic Scientists, 7(12):383–384, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:CDNa

[Ano51e] Anonymous. Civil defense news. Bulletin of the Atomic Scientists, 7(4):127, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:CDNb

[Ano51f] Anonymous. Civil defense news. Bulletin of the Atomic Scientists, 7(5):158–160, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:CDNc

[Ano51g] Anonymous. Civil defense news. Bulletin of the Atomic Scientists, 7(9):285–286, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:CE

[Ano51h] Anonymous. Comfort to the enemy. Bulletin of the Atomic Scientists, 7(3):76, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:CMF

[Ano51i] Anonymous. Council meeting of the Federation of American Scientists. Bulletin of the Atomic Scientists, 7(6):182–183, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:CL

[Ano51j] Anonymous. A curtain is lifted. Bulletin of the Atomic Scientists, 7(9):286, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:ECD

[Ano51k] Anonymous. Editorial: Civil defense and the Bulletin. Bulletin of the Atomic Scientists, 7(10):319, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:EPC

[Ano51l] Anonymous. The electric power company and the AEC: Report of an Advisory Committee on Cooperation between the Electric Power Industry and the Atomic Energy Commission. Bulletin of the Atomic Scientists, 7(11):332–336, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:ET

[Ano51m] Anonymous. The Eniwetok tests. Bulletin of the Atomic Scientists, 7(7–8):239, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:FCD

[Ano51n] Anonymous. Federal Civil Defense Act. Bulletin of the Atomic Scientists, 7(2):59–62, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:HHG

[Ano51o] Anonymous. H. H. Goldsmith Award. Bulletin of the Atomic Scientists, 7(5):131, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:IDN

[Ano51p] Anonymous. Industrial dispersion as a national policy. Bulletin of the Atomic Scientists, 7(9):265–269, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:SBI

[Ano51q] Anonymous. Is “Strategic bombing” ineffective? Bulletin of the Atomic Scientists, 7(9):288, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:JCR

[Ano51r] Anonymous. The Joint Committee review the U.S. Atomic Energy Project. Bulletin of the Atomic Scientists, 7(11):342–344, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:MM

[Ano51s] Anonymous. Mobilization of manpower. Bulletin of the Atomic Scientists, 7(3):88, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:NCW

[Ano51t] Anonymous. A news chronology: Weeks of history and mystery. Bulletin of the Atomic Scientists, 7(10):302–304, 320, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:NS

[Ano51u] Anonymous. The NSF story. Bulletin of the Atomic Scientists, 7(10):291, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:NDP

[Ano51v] Anonymous. NSRB dispersion program. Bulletin of the Atomic Scientists, 7(9):263, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:PS

[Ano51w] Anonymous. Plea for secrecy. Bulletin of the Atomic Scientists, 7(9):277, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:QC

[Ano51x] Anonymous. Quotes without comment. Bulletin of the Atomic Scientists, 7(11):349, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:RBA

[Ano51y] Anonymous. Report of British Atomic Scientists annual meeting. Bulletin of the Atomic Scientists, 7(11):345, November 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:SOS

[Ano51z] Anonymous. Senators oppose secrecy. Bulletin of the Atomic Scientists, 7(11):352, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:UCT

[Ano51-27] Anonymous. ”The universal casualty tag … ”. Bulletin of the Atomic Scientists, 7(9):285, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:UNA

[Ano51-28] Anonymous. United Nations atomic energy news. Bulletin of the Atomic Scientists, 7(1):28–29, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1951:VLB

[Ano51-29] Anonymous. V: Lord Boyd Orr on the responsibilities of scientists. Bulletin of the Atomic Scientists, 7(12):379, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:AGA

[Ano52a] Anonymous. AEC General Advisory Committee issues report. Bulletin of the Atomic Scientists, 8(8):288, November 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:ACU

[Ano52b] Anonymous. Announcement by the Chairman, United States Atomic Energy Commission (November 16, 1952). Bulletin of the Atomic Scientists, 8(9):298, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:BWF

[Ano52c] Anonymous. Biological warfare: FAS asks clarification of U.S. policies on germ warfare. Bulletin of the Atomic Scientists, 8 (5):130, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:BR

[Ano52d] Anonymous. Books received. Bulletin of the Atomic Scientists, 8(5):165, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:BAS

[Ano52e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume VIII, 1952. Bulletin of the Atomic Scientists, 8(9):315–316, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:CSP

[Ano52f] Anonymous. Correction: Soviet psychology and psychiatry. Bulletin of the Atomic Scientists, 8(5):162, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lon52a].

Anonymous:1952:DFM

[Ano52g] Anonymous. A declaration by French members of the teaching profession in view of the menace of a new World War. Bulletin of the Atomic Scientists, 8(7):243–244, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Anonymous:1952:EMH

[Ano52h] Anonymous. Excerpts from MacArthur hearings, May 4, 1951. Bulletin of the Atomic Scientists, 8(1):14–15, January 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:ESB

[Ano52i] Anonymous. Eyewitness stories of the bomb test. Bulletin of the Atomic Scientists, 8(9):300, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:FSR

[Ano52j] Anonymous. Free speech for the “Red Dean”. Bulletin of the Atomic Scientists, 8(6):181, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From a debate in the British House of Commons.

Anonymous:1952:GRP

[Ano52k] Anonymous. The government reactor program: An interview with Lawrence Halfstad: [Director of the AEC’s Division of Reactor Development]. Bulletin of the Atomic Scientists, 8(8):283– 286, 292, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:HNS

[Ano52l] Anonymous. Hearings of the National Science Foundation budget. Bulletin of the Atomic Scientists, 8(4):104–110, April 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:MA

[Ano52m] Anonymous. The McCarran Act. Bulletin of the Atomic Scientists, 8(7):257–258, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Anonymous:1952:NNa

[Ano52n] Anonymous. News and notes. Bulletin of the Atomic Scientists, 8(6):207–208, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:NNb

[Ano52o] Anonymous. News and notes. Bulletin of the Atomic Scientists, 8 (8):291–292, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:NNc

[Ano52p] Anonymous. News and notes. Bulletin of the Atomic Scientists, 8 (9):309–312, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:NMW

[Ano52q] Anonymous. A note on Mr. Wallace’s article. Bulletin of the Atomic Scientists, 8(2):37–38, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:OBA

[Ano52r] Anonymous. Oppenheimer and Bush appointed to new Advisory Group on Disarmament. Bulletin of the Atomic Scientists, 8(5): 133, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:PTE

[Ano52s] Anonymous. President Truman’s endorsement of National Science Foundation — January 15, 1952. Bulletin of the Atomic Scientists, 8(4):105, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:PIA

[Ano52t] Anonymous. Private industry and atomic power. Bulletin of the Atomic Scientists, 8(8):275, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:QC

[Ano52u] Anonymous. Quotes without comments. Bulletin of the Atomic Scientists, 8(2):64, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:SUb

[Ano52v] Anonymous. Science and UNESCO. Bulletin of the Atomic Scientists, 8(9):308, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:SP

[Ano52w] Anonymous. “Science for peace”. Bulletin of the Atomic Scientists, 8(4):128, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:SUa

[Ano52x] Anonymous. “Science is universal”. Bulletin of the Atomic Scientists, 8(2):38–39, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:SDJ

[Ano52y] Anonymous. Statement by Dr. Jules Halpern, Chairman, Federation of American Scientists. Bulletin of the Atomic Scientists, 8(9):299, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:SB

[Ano52z] Anonymous. Strangely beautiful. Bulletin of the Atomic Scientists, 8(6):202–203, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Report of Nevada bomb test.

Anonymous:1952:UNA

[Ano52-27] Anonymous. United Nations atomic energy news; U.S. proposals for world arms census. Bulletin of the Atomic Scientists, 8(5): 135–138, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1952:WGR

[Ano52-28] Anonymous. Walter Gellhorn receives the Hyman Goldsmith Memorial Award. Bulletin of the Atomic Scientists, 8(3):76, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:AEPa

[Ano53a] Anonymous. Atomic energy and private enterprise: Joint Committee hearings. Bulletin of the Atomic Scientists, 9(8):309–314, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:AEPb

[Ano53b] Anonymous. Atomic energy and private enterprise: Joint Committee hearings. Bulletin of the Atomic Scientists, 9(9):341–344, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:AEPc

[Ano53c] Anonymous. Atomic energy and private enterprise: Joint Committee hearings. Bulletin of the Atomic Scientists, 9(10):380–382, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:APP

[Ano53d] Anonymous. Atomic power and private enterprise: a summary of the Joint Committee report. Bulletin of the Atomic Scientists, 9 (4):135–140, 144, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BNB

[Ano53e] Anonymous. Book note: The Radiation Laboratory Series Index. Bulletin of the Atomic Scientists, 9(8):320, October 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BRa

[Ano53f] Anonymous. Books received. Bulletin of the Atomic Scientists, 9 (1):32, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BRb

[Ano53g] Anonymous. Books received. Bulletin of the Atomic Scientists, 9(3):92, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BRc

[Ano53h] Anonymous. Books received. Bulletin of the Atomic Scientists, 9(9):350, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BRd

[Ano53i] Anonymous. Books received. Bulletin of the Atomic Scientists, 9(10):390, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BAS

[Ano53j] Anonymous. Bulletin of the Atomic Scientists: Index to Volume IX, 1953. Bulletin of the Atomic Scientists, 9(10):391–392, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:BDA

[Ano53k] Anonymous. British debates for atomic power. Bulletin of the Atomic Scientists, 9(4):141–144, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:CCS

[Ano53l] Anonymous. Current comment: Science and government. Bulletin of the Atomic Scientists, 9(4):151, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:CCG

[Ano53m] Anonymous. Current comment: The great debate. Bulletin of the Atomic Scientists, 9(9):348–349, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:DCC

[Ano53n] Anonymous. Defense considerations in city planning: Statement by the American Institute of Planners. Bulletin of the Atomic Scientists, 9(7):268, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:ENS

[Ano53o] Anonymous. Editor’s note: [Soviet Government’s announcement of the first successful hydrogen bomb test]. Bulletin of the Atomic Scientists, 9(7):235, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:E

[Ano53p] Anonymous. Erratum. Bulletin of the Atomic Scientists, 9(9): ??, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See page 295 of October 1953 issue.

Anonymous:1953:EPN

[Ano53q] Anonymous. Erratum: Page 351 of November 1953 issue of the Bulletin. Bulletin of the Atomic Scientists, 9(10):379, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:FFSa

[Ano53r] Anonymous. Federal funds for science. Bulletin of the Atomic Scientists, 9(9):345–347, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:FFSb

[Ano53s] Anonymous. Federal funds for science. Bulletin of the Atomic Scientists, 9(10):389, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:FOB

[Ano53t] Anonymous. Functions and operations of the Bureau of Standards: a summary of the Kelly Report. Bulletin of the Atomic Scientists, 9(10):376–377, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:HBS

[Ano53u] Anonymous. The H-bomb story. Bulletin of the Atomic Scientists, 9(1):6, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:IGV

[Ano53v] Anonymous. An interview with Governor Val Peterson. Bulletin of the Atomic Scientists, 9(7):237–242, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:JRE

[Ano53w] Anonymous. Joint resolution to establish a Joint Committee on Science. Bulletin of the Atomic Scientists, 9(4):108, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NBS

[Ano53x] Anonymous. National Bureau of Standards: Astin–Weeks case. Bulletin of the Atomic Scientists, 9(4):103–104, 146, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NGE

[Ano53y] Anonymous. The new Government Employee Security Program. Bulletin of the Atomic Scientists, 9(5):175, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNa

[Ano53z] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(1):29–31, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNb

[Ano53-27] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(2):62–64, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNc

[Ano53-28] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(3):90–91, 94, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNd

[Ano53-29] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(4):145, 150, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNe

[Ano53-30] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(5):195–196, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNf

[Ano53-31] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(6):229–231, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNg

[Ano53-32] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9 (7):285–288, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNh

[Ano53-33] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9(8):318–319, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNi

[Ano53-34] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9 (9):351–352, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:NNj

[Ano53-35] Anonymous. News and notes. Bulletin of the Atomic Scientists, 9 (10):383–387, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:OA

[Ano53-36] Anonymous. Our apologies. Bulletin of the Atomic Scientists, 9 (5):190, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:RRU

[Ano53-37] Anonymous. The rights and responsibilities of universities. Bulletin of the Atomic Scientists, 9(5):187, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:RTT

[Ano53-38] Anonymous. Russian thermonuclear test. Bulletin of the Atomic Scientists, 9(7):236, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:SP

[Ano53-39] Anonymous. Scientists and the Press. Bulletin of the Atomic Scientists, 9(9):328, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:SDT

[Ano53-40] Anonymous. Scientists defend themselves on loyalty charges. Bulletin of the Atomic Scientists, 9(1):2–3, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:SBM

[Ano53-41] Anonymous. The Soviet bombs: Mr. Truman’s doubts. Bulletin of the Atomic Scientists, 9(2):43, 45, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:SFE

[Ano53-42] Anonymous. Statement by FAS Executive Committee. Bulletin of the Atomic Scientists, 9(10):377, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:SAE

[Ano53-43] Anonymous. Symposium: Atomic energy and private enterprise: Part I. Bulletin of the Atomic Scientists, 9(4):109, May 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:OR

[Ano53-44] Anonymous. To our readers. Bulletin of the Atomic Scientists, 9(3):93, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:UDT

[Ano53-45] Anonymous. Unesco defended but technical assistance curtailed. Bulletin of the Atomic Scientists, 9(9):349–350, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:UGW

[Ano53-46] Anonymous. Use of gas in warfare; atom bombs and tornados; visa policy and the duty of scientists. Bulletin of the Atomic Scientists, 9(6):227–228, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:WAC

[Ano53-47] Anonymous. Weeks–Astin controversy. Bulletin of the Atomic Scientists, 9(3):94, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1953:WWS

[Ano53-48] Anonymous. “Whom we shall welcome”. Bulletin of the Atomic Scientists, 9(1):17–21, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:AAP

[Ano54a] Anonymous. Against atomic panic. Bulletin of the Atomic Scientists, 10(6):229, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:ASD

[Ano54b] Anonymous. American scientist declared security risk. Bulletin of the Atomic Scientists, 10(6):243, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:ASB

[Ano54c] Anonymous. “Atomic scientists’ Bulletin does science disservice”. Bulletin of the Atomic Scientists, 10(6):213, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:BRd

[Ano54d] Anonymous. Books received. Bulletin of the Atomic Scientists, 10(4):137, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:BRf

[Ano54e] Anonymous. Books received. Bulletin of the Atomic Scientists, 10(6):230, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:BAS

[Ano54f] Anonymous. Bulletin of the Atomic Scientists: Index to Volume X, 1954. Bulletin of the Atomic Scientists, 10(10):399–400, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:BRe

[Ano54g] Anonymous. The British reaction. Bulletin of the Atomic Scientists, 10(5):166, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:CAV

[Ano54h] Anonymous. Congressional action on visa problems: Joint resolution. Bulletin of the Atomic Scientists, 10(3):70–71, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:CCM

[Ano54i] Anonymous. Current comment: The making of the H-bomb. Bulletin of the Atomic Scientists, 10(6):228–229, June 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from an article of the same title in Time on 12 April 1954.

Anonymous:1954:DAI

[Ano54j] Anonymous. Dispersal of American industry. Bulletin of the Atomic Scientists, 10(6):196, 205, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:ERB

[Ano54k] Anonymous. Erratum: Review of The Making of a Scientist. Bulletin of the Atomic Scientists, 10(9):344, November 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch54a].

Anonymous:1954:FNa

[Ano54l] Anonymous. FAS news. Bulletin of the Atomic Scientists, 10(1): 27, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:FNb

[Ano54m] Anonymous. FAS news. Bulletin of the Atomic Scientists, 10(2): 64, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:FMI

[Ano54n] Anonymous. The Fort Monmouth investigations. Bulletin of the Atomic Scientists, 10(6):225–226, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:FAE

[Ano54o] Anonymous. Franchise for atomic energy. Bulletin of the Atomic Scientists, 10(3):100–102, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:LCC

[Ano54p] Anonymous. L’Affaire Curie: The chemist and the Communist. Bulletin of the Atomic Scientists, 10(6):211–, June 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the April 1954 issue of The Chemical Bulletin. See also [Hai54, Eme54].

Anonymous:1954:MAU

[Ano54q] Anonymous. Major activities in the U.S. atomic energy program. I. The reactor program. Bulletin of the Atomic Scientists, 10(9): 360–362, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:NN

[Ano54r] Anonymous. News and notes. Bulletin of the Atomic Scientists, 10(1):32, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:NR

[Ano54s] Anonymous. News roundup. Bulletin of the Atomic Scientists, 10 (9):365–368, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:NRA

[Ano54t] Anonymous. News roundup: Atomic power; the arms race; atomic energy abroad; American science; world science; USAEC news; items. Bulletin of the Atomic Scientists, 10(8):332–334, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:NRN

[Ano54u] Anonymous. News roundup: The new Atomic Energy Act; disarmament; peaceful uses; atomic power here and abroad; the arms race; U.S. Atomic Energy Commission; nuclear developments abroad. Bulletin of the Atomic Scientists, 10(7):301–304, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:OH

[Ano54v] Anonymous. News roundup: The Oppenheimer hearings; the arms race; the peaceful atom; continental defense. Bulletin of the Atomic Scientists, 10(6):234–237, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:OL

[Ano54w] Anonymous. [Oppenheimer’s loyalty]. Bulletin of the Atomic Scientists, 10(5):185, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:PEC

[Ano54x] Anonymous. Planning an effective civilian defense. Bulletin of the Atomic Scientists, 10(7):291–293, 295, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:RBE

[Ano54y] Anonymous. Reaction in the Chemical and Engineering News. Bulletin of the Atomic Scientists, 10(6):214, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:ROT

[Ano54z] Anonymous. Read the Oppenheimer transcript. Bulletin of the Atomic Scientists, 10(7):258, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:SUS

[Ano54-27] Anonymous. Scientists and United States visas. Bulletin of the Atomic Scientists, 10(4):145, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:SEC

[Ano54-28] Anonymous. Scientists express confidence in Oppenheimer. Bulletin of the Atomic Scientists, 10(7):283, 286, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:SNE

[Ano54-29] Anonymous. Soviet note on Eisenhower proposal. Bulletin of the Atomic Scientists, 10(2):46–47, 54, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:SPL

[Ano54-30] Anonymous. Soviet peoples learn facts of atomic age. Bulletin of the Atomic Scientists, 10(5):167–168, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1954:STA

[Ano54-31] Anonymous. Symposium: Technical assistance programs. Part I. Bulletin of the Atomic Scientists, 10(3):72, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:ACS

[Ano55a] Anonymous. AEC criteria for security clearance. Bulletin of the Atomic Scientists, 11(4):159–160, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Anonymous:1955:BAI

[Ano55b] Anonymous. A bill to amend the Immigration and Nationality Act. Bulletin of the Atomic Scientists, 11(5):199, May 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:BBF

[Ano55c] Anonymous. Books: Psychological First Aid in Community Disasters. Bulletin of the Atomic Scientists, 11(2):64, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:BAS

[Ano55d] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XI, 1955. Bulletin of the Atomic Scientists, 11(10):383–384, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:CCK

[Ano55e] Anonymous. Candor in Congress: The Kefauver Hearings. Bulletin of the Atomic Scientists, 11(5):181–184, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:CRP

[Ano55f] Anonymous. Condon resigns post. Bulletin of the Atomic Scientists, 11(1):40, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:DNP

[Ano55g] Anonymous. Development of a nuclear-powered merchant ship. Bulletin of the Atomic Scientists, 11(6):229–230, June 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:ES

[Ano55h] Anonymous. The Earth satellite. Bulletin of the Atomic Scientists, 11(7):243, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:ECS

[Ano55i] Anonymous. Edward Condon’s security clearance: Statement by Corning Glass Works. Bulletin of the Atomic Scientists, 11(2): 66, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:ELS

[Ano55j] Anonymous. Erratum: “Loyalty–Security Measures and Employment Opportunities”. Bulletin of the Atomic Scientists, 11(6): 229, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro55c].

Anonymous:1955:FP

[Ano55k] Anonymous. Fermi Prize. Bulletin of the Atomic Scientists, 11(2): 70, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The announcement begins: “The first AEC award of $25,000 for especially meritorious contribution to the development of atomic energy, provided for in the Atomic Energy Act of 1954, was given to Enrico Fermi, November 16, 1954, twelve days before his death.” It end with “Subsequent awards will be known as the “Fermi Prize”.

Anonymous:1955:FIC

[Ano55l] Anonymous. First International Conference on Atomic Energy, Geneva, August 8–20, 1955. Bulletin of the Atomic Scientists, 11(8):274, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:FMO

[Ano55m] Anonymous. Fort Monmouth one year later: Scientists’ Committee on Loyalty and Security. Bulletin of the Atomic Scientists, 11(4):148–150, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Anonymous:1955:GGT

[Ano55n] Anonymous. Glossary of genetics terms. Bulletin of the Atomic Scientists, 11(9):317, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:NVF

[Ano55o] Anonymous. A nonimmigration visa form. Bulletin of the Atomic Scientists, 11(10):374–375, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:NCS

[Ano55p] Anonymous. [Note on Conference on Science and Freedom, Hamburg, Germany, July 23–26, 1953]. Bulletin of the Atomic Scientists, 11(7):268, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shi54a].

Anonymous:1955:PES

[Ano55q] Anonymous. President Eisenhower speaks on disarmament: Big Four Conference, Geneva, July 21, 1955. Bulletin of the Atomic Scientists, 11(7):263–264, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:PUN

[Ano55r] Anonymous. Proposal for a United Nations Commission to study the problem of H-bomb tests. Bulletin of the Atomic Scientists, 11(5):185–186, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:RDU

[Ano55s] Anonymous. Recent developments in U.S. uranium mining. Bulletin of the Atomic Scientists, 11(3):91, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:RAP

[Ano55t] Anonymous. Resolution on the atoms-for-peace plan. Bulletin of the Atomic Scientists, 11(1):27, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:RSM

[Ano55u] Anonymous. Roentgens, sickness, and mortality. Bulletin of the Atomic Scientists, 11(2):48, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:SCS

[Ano55v] Anonymous. The Scientists’ Committee on Security, Inc. Bulletin of the Atomic Scientists, 11(4):147, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Anonymous:1955:SRG

[Ano55w] Anonymous. Security requirements for government employment: Executive Order 10450. Bulletin of the Atomic Scientists, 11(4): 156–158, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Anonymous:1955:SRV

[Ano55x] Anonymous. Some recent visa experiences. Bulletin of the Atomic Scientists, 11(10):368–371, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1955:UND

[Ano55y] Anonymous. United Nations Disarmament Commission Subcommittee: Summary of most recent proposals, London, February 25–May 10, 1955. Bulletin of the Atomic Scientists, 11(7):261– 262, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:A

[Ano56a] Anonymous. Addendum. Bulletin of the Atomic Scientists, 12 (5):175, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol56].

Anonymous:1956:BBT

[Ano56b] Anonymous. Books: Experiment and Theory in Physics, by Max Born. Bulletin of the Atomic Scientists, 12(7):280, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:BBR

[Ano56c] Anonymous. Books: Shelter from Radioactive Fall-out, and Soviet Professional Manpower, by Nicholas DeWitt. Bulletin of the Atomic Scientists, 12(1):25, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:BBM

[Ano56d] Anonymous. Books: The Men Behind the Space Rockets, by Heinz Gartmann. Bulletin of the Atomic Scientists, 12(9):349, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:BAS

[Ano56e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XII, 1956. Bulletin of the Atomic Scientists, 12(10):383–384, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:CTR

[Ano56f] Anonymous. Correction: Today’s revolution. Bulletin of the Atomic Scientists, 12(10):380, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei56a].

Anonymous:1956:MHT

[Ano56g] Anonymous. The memoirs of Harry S. Truman: Atom problems arise. Bulletin of the Atomic Scientists, 12(4):132–134, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:NMT

[Ano56h] Anonymous. A Navy medical team studies fall-out effects. Bulletin of the Atomic Scientists, 12(2):58–59, February 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the Office of Naval Research publication, Research Reviews. The article discusses fall-out on four inhabited islands of the Marshall Islands.

Anonymous:1956:NWT

[Ano56i] Anonymous. The nuclear weapons test ban. Bulletin of the Atomic Scientists, 12(7):268, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:RPP

[Ano56j] Anonymous. Report of the Panel on the Peaceful Uses of Atomic Energy. Bulletin of the Atomic Scientists, 12(3):84–89, 96, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:SCS

[Ano56k] Anonymous. The Scientists’ Committee on Security. Bulletin of the Atomic Scientists, 12(8):321, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:STL

[Ano56l] Anonymous. Summary of testimony of Linus Pauling. Bulletin of the Atomic Scientists, 12(1):28, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:TBD

[Ano56m] Anonymous. Test ban dialogue (up to October 15 [1956]). Bulletin of the Atomic Scientists, 12(9):324, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1956:UAC

[Ano56n] Anonymous. U.N. appoints committee to study effects of ionizing radiation. Bulletin of the Atomic Scientists, 12(1):13, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:AAA

[Ano57a] Anonymous. American Association for the Advancement of Science: Interim Committee Report. Bulletin of the Atomic Scientists, 13(3):81–85, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:BBH

[Ano57b] Anonymous. Books: Calder Hall. The Story of Britain’s First Atomic Power Station, by Kenneth Jay. Bulletin of the Atomic Scientists, 13(2):72, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:BBA

[Ano57c] Anonymous. Books: International Atomic Policy During a Decade, by Admiral Elis Biorklund. Bulletin of the Atomic Scientists, 13(1):35, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:BBE

[Ano57d] Anonymous. Books: Nuclear Energy in Western Europe and United States Policy, by Klaus E. Knorr. Bulletin of the Atomic Scientists, 13(1):35, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:BAS

[Ano57e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XIII, 1957. Bulletin of the Atomic Scientists, 13(10):375–376, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:DJG

[Ano57f] Anonymous. Dr. Joseph G. Hamilton (1907–1956). Bulletin of the Atomic Scientists, 13(4):149, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:FC

[Ano57g] Anonymous. Fair comment? Bulletin of the Atomic Scientists, 13(7):250, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from an editorial What Goes on? in the Chicago Daily Tribune, 16 August 1957, objecting to the presence of a ‘Red’ Chinese scientist [Pei-Yuan Chou] at the Pugwash Conference, and charging Bertrand Russell and Albert Einstein with having “a certain fondness for communism”.

Anonymous:1957:JPA

[Ano57h] Anonymous. Japanese physicists address their British colleagues. Bulletin of the Atomic Scientists, 13(5):178, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The appeal to end nuclear testing is signed by over 350 Japanese physicists.

Anonymous:1957:PCS

[Ano57i] Anonymous. The Pugwash Conference: Statement. Bulletin of the Atomic Scientists, 13(7):249–250, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). “The release of the following statement with the three committee reports was approved by all the above-listed participants, with the exception of Dr. Foster, who objected to the Report of Committee II and Dr. Szilard, who objected to those of Committees I and II.”.

Anonymous:1957:SIG

[Ano57j] Anonymous. Statement issued at the general convention of the Elementary Particle Theory Group, October 30, 1956, Kyoto, Japan. Bulletin of the Atomic Scientists, 13(2):73, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:SGS

[Ano57k] Anonymous. Statement of a group of Soviet scientists concerning the quest of prohibition of atomic weapons and their tests. Bulletin of the Atomic Scientists, 13(9):316–317, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1957:MM

[Ano57l] Anonymous. “The Moscow Missile”. Bulletin of the Atomic Scientists, 13(8):301, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from an editorial in the New York Times on 28 August 1957.

Anonymous:1958:AFH

[Ano58a] Anonymous. An appeal — not for funds, but for history. Bulletin of the Atomic Scientists, 14(8):312, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:BAS

[Ano58b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XIV, 1958. Bulletin of the Atomic Scientists, 14(10):431–432, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:BPG

[Ano58c] Anonymous. British Parliamentary Group for World Government. Bulletin of the Atomic Scientists, 14(6):232–233, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:EPE

[Ano58d] Anonymous. Excerpts from President Eisenhower’s State of the Union message, January 9, 1958. Bulletin of the Atomic Scientists, 14(2):90, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:ECT

[Ano58e] Anonymous. Experts’ conclusions on test detection. Bulletin of the Atomic Scientists, 14(8):329–330, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:G

[Ano58f] Anonymous. Glossary. Bulletin of the Atomic Scientists, 14(1): 65, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:LBQ

[Ano58g] Anonymous. Lac Beauport, Quebec, March 31–April 11, 1958: Statement from the Second Pugwash Conference. Bulletin of the Atomic Scientists, 14(5):195, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:PIA

[Ano58h] Anonymous. The press interviews AEC Commissioner Libby. Bulletin of the Atomic Scientists, 14(5):190–192, May 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:RSW

[Ano58i] Anonymous. A report to the Secretary of War — June 1945. Bulletin of the Atomic Scientists, 14(8):302, October 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from the Franck Report [FHN^{+}46].

Anonymous:1958:SPO

[Ano58j] Anonymous. Soviet proposal on outer space. Bulletin of the Atomic Scientists, 14(5):185–186, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1958:VDT

[Ano58k] Anonymous. The Vienna declaration: Third Pugwash Conference of Nuclear Scientists: Kitzbühel-Vienna, Austria, September 14– 21, 1958. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1958. 11 pp. LCCN ????

Anonymous:1958:VDS

[Ano58l] Anonymous. Vienna Declaration: Statement of the conference on “The Dangers of the Atomic Age and What Scientists Can Do About Them” (Third Pugwash Conference), held at Kitzbuhel, Austria on September 14–19, and at Vienna on September 20, 1958. Bulletin of the Atomic Scientists, 14(9):341–344, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See note [Rab59b].

Anonymous:1959:WLF

[Ano59a] Anonymous. Are we letting the future go by default? Bulletin of the Atomic Scientists, 15(1):13, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:BAS

[Ano59b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XV, 1959. Bulletin of the Atomic Scientists, 15(10):431–432, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:DPS

[Ano59c] Anonymous. Democratic planning for science. Bulletin of the Atomic Scientists, 15(10):412, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:FPC

[Ano59d] Anonymous. Fourth Pugwash Conference: Baden, near Vienna, June 25–July 4, 1959. Bulletin of the Atomic Scientists, 15(7): 310, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:THS

[Ano59e] Anonymous. From two historic speeches [N.S. Khrushchev]. Bulletin of the Atomic Scientists, 15(9):366, November 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:GG

[Ano59f] Anonymous. Geological glossary. Bulletin of the Atomic Scientists, 15(6):273, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:N

[Ano59g] Anonymous. Notes. Bulletin of the Atomic Scientists, 15(8):350, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:NSP

[Ano59h] Anonymous. Notes: Specialized pamphlets. Bulletin of the Atomic Scientists, 15(4):173, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:RGC

[Ano59i] Anonymous. Roadblocks at Geneva: Conference on the discontinuance of nuclear weapons tests. Bulletin of the Atomic Scientists, 15(3):137–138, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1959:USN

[Ano59j] Anonymous. U.S. statement on nuclear tests. Bulletin of the Atomic Scientists, 15(2):96, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:ACT

[Ano60a] Anonymous. Annex I: Conclusion of Technical Working Group II regarding possible improvements of techniques and instrumentation. Bulletin of the Atomic Scientists, 16(2):38–39, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:AIS

[Ano60b] Anonymous. Annex II: Statement by Soviet experts. Bulletin of the Atomic Scientists, 16(2):39–43, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:AIRa

[Ano60c] Anonymous. Annex III: Report by the United Kingdom delegation from the Proceedings of Technical Working Group II. Bulletin of the Atomic Scientists, 16(2):44–45, February 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:AIRb

[Ano60d] Anonymous. Annex IV: Report of the delegation of the United States concerning Technical Working Group II of the Conference on the Discontinuance of Nuclear Weapons Tests. Bulletin of the Atomic Scientists, 16(2):45–47, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:ACD

[Ano60e] Anonymous. Appendix: Comments by the delegation of the United States. Bulletin of the Atomic Scientists, 16(2):48, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:BBP

[Ano60f] Anonymous. Books: Science and Public Policy, by Dael Wolfle; The N th Country Problem and Arms Control; Towards a New Energy Pattern in Europe; The Industrial Challenge of Nuclear Energy; An Atomic Development Plan for the State of New York; Nuclear Physics and Atomic Energy by Georges J. Béné, R. Beeler, and M. Golub. Bulletin of the Atomic Scientists, 16 (8):341, 344, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:BAS

[Ano60g] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XVI, 1960. Bulletin of the Atomic Scientists, 16(10):421–423, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:BRB

[Ano60h] Anonymous. Briefs [recent books]. Bulletin of the Atomic Scientists, 16(9):371, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:BEN

[Ano60i] Anonymous. [British experiments with nerve gases]. Bulletin of the Atomic Scientists, 16(6):251, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:DCS

[Ano60j] Anonymous. Disarmament — comparison of Soviet and Western proposals. Bulletin of the Atomic Scientists, 16(8):336–339, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:FDN

[Ano60k] Anonymous. The future development of nuclear weapons. Bulletin of the Atomic Scientists, 16(5):167, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Dys60].

Anonymous:1960:MR

[Ano60l] Anonymous. The McKinney Report. Bulletin of the Atomic Scientists, 16(9):354, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:PTS

[Ano60m] Anonymous. Peace, testing, and space. Bulletin of the Atomic Scientists, 16(6):219, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:SPPb

[Ano60n] Anonymous. Science and party politics: Democrats speak. Bulletin of the Atomic Scientists, 16(7):264, September 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:SPPa

[Ano60o] Anonymous. Science and party politics: Nuclear testing. Bulletin of the Atomic Scientists, 16(3):109, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:SPPc

[Ano60p] Anonymous. Science and party politics: Republicans speak. Bulletin of the Atomic Scientists, 16(7):265, September 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:SPPd

[Ano60q] Anonymous. Science and party politics: Science and a better America. Bulletin of the Atomic Scientists, 16(8):340, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:SPPe

[Ano60r] Anonymous. Soviet position on peaceful detonations: Geneva test ban negotiations. Bulletin of the Atomic Scientists, 16(9): 369, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:SP

[Ano60s] Anonymous. Specialized pamphlets. Bulletin of the Atomic Scientists, 16(10):416–417, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:WP

[Ano60t] Anonymous. What is Pugwash? Bulletin of the Atomic Scientists, 16(8):304, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:A

[Ano61a] Anonymous. Authors. Bulletin of the Atomic Scientists, 17(10): 449, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:BBa

[Ano61b] Anonymous. Book briefs. Bulletin of the Atomic Scientists, 17 (4):163, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:BBb

[Ano61c] Anonymous. Book briefs. Bulletin of the Atomic Scientists, 17 (10):437–438, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:BN

[Ano61d] Anonymous. Book notes. Bulletin of the Atomic Scientists, 17(7): 292, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:BAS

[Ano61e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XVII, 1961. Bulletin of the Atomic Scientists, 17(10):446–448, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:BRBa

[Ano61f] Anonymous. Briefs [recent books]. Bulletin of the Atomic Scientists, 17(1):37, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:BRBb

[Ano61g] Anonymous. Briefs [recent books]. Bulletin of the Atomic Scientists, 17(3):116–117, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:C

[Ano61h] Anonymous. Correction. Bulletin of the Atomic Scientists, 17(2): 81, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [All60g].

Anonymous:1961:DAC

[Ano61i] Anonymous. Disarmament and arms control: Statement from the Sixth Pugwash Conference, Moscow, November 27–December 5, 1960. Bulletin of the Atomic Scientists, 17(2):74–75, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:EC

[Ano61j] Anonymous. Echo from the campaign. Bulletin of the Atomic Scientists, 17(2):54, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:EAS

[Ano61k] Anonymous. Excerpts from AEC statement on nuclear blast effects. Bulletin of the Atomic Scientists, 17(10):445, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:G

[Ano61l] Anonymous. Glossary. Bulletin of the Atomic Scientists, 17(5–6): 257, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:GR

[Ano61m] Anonymous. Grant revoked. Bulletin of the Atomic Scientists, 17(7):299, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:MS

[Ano61n] Anonymous. Man in space. Bulletin of the Atomic Scientists, 17 (10):427–428, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:NRN

[Ano61o] Anonymous. News and reports: Neutron bomb: How, why, when? Bulletin of the Atomic Scientists, 17(7):297, September 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:NRT

[Ano61p] Anonymous. News and reports: Test resumption and fallout. Bulletin of the Atomic Scientists, 17(9):394–395, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:RBS

[Ano61q] Anonymous. Recent books on space science. Bulletin of the Atomic Scientists, 17(5–6):246, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:SGR

[Ano61r] Anonymous. Science and government: Report. Bulletin of the Atomic Scientists, 17(7):298, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:SD

[Ano61s] Anonymous. Space developments. Bulletin of the Atomic Scientists, 17(10):443–444, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:SN

[Ano61t] Anonymous. Space news. Bulletin of the Atomic Scientists, 17(5– 6):247–249, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:SP

[Ano61u] Anonymous. Specialized pamphlets. Bulletin of the Atomic Scientists, 17(2):77, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:TF

[Ano61v] Anonymous. Tests and fallout. Bulletin of the Atomic Scientists, 17(10):442–443, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:WWa

[Ano61w] Anonymous. Who’s who. Bulletin of the Atomic Scientists, 17(8): 352, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1961:WWb

[Ano61x] Anonymous. Who’s who. Bulletin of the Atomic Scientists, 17(9): 400, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Aa

[Ano62a] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(1): 48, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ab

[Ano62b] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(2): 41, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ac

[Ano62c] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(3):49, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ad

[Ano62d] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(4):48, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ae

[Ano62e] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(5):48, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Af

[Ano62f] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(6):48, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ag

[Ano62g] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(7): 48, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ah

[Ano62h] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(8): 48, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Ai

[Ano62i] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(9): 48, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:Al

[Ano62j] Anonymous. Authors. Bulletin of the Atomic Scientists, 18(10): 45, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:BR

[Ano62k] Anonymous. Book reviews. Bulletin of the Atomic Scientists, 18 (4):42, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:BBP

[Ano62l] Anonymous. Books: Population Perspectives, by Philip M. Hauser. Bulletin of the Atomic Scientists, 18(5):40, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:BAS

[Ano62m] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XVIII, 1962. Bulletin of the Atomic Scientists, 18(10):46–48, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:BRB

[Ano62n] Anonymous. Briefs [recent books]. Bulletin of the Atomic Scientists, 18(1):41, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:CD

[Ano62o] Anonymous. Civil defense. Bulletin of the Atomic Scientists, 18(1):45–46, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:CDA

[Ano62p] Anonymous. Civil defense against nuclear attack: A selected bibliography. Bulletin of the Atomic Scientists, 18(7):33–34, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:CDT

[Ano62q] Anonymous. Comparison of disarmament treaties. Bulletin of the Atomic Scientists, 18(7):8–9, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:FM

[Ano62r] Anonymous. Fallout measurements. Bulletin of the Atomic Scientists, 18(1):44–45, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:IN

[Ano62s] Anonymous. International news. Bulletin of the Atomic Scientists, 18(1):47–48, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:NN

[Ano62t] Anonymous. Names in the news. Bulletin of the Atomic Scientists, 18(5):48, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:OSC

[Ano62u] Anonymous. Official statement of COSWA VII. Bulletin of the Atomic Scientists, 18(1):25–30, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:OLN

[Ano62v] Anonymous. Open letter on nuclear policy. Bulletin of the Atomic Scientists, 18(9):41–42, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:OLP

[Ano62w] Anonymous. An open letter to President Kennedy. Bulletin of the Atomic Scientists, 18(2):28–29, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter in opposition to civil defense signed by about 200 Boston-area university staff members. See response [Ano62y].

Anonymous:1962:PCS

[Ano62x] Anonymous. Pugwash–COSWA statement. Bulletin of the Atomic Scientists, 18(9):40, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:RL

[Ano62y] Anonymous. Reactions to letter. Bulletin of the Atomic Scientists, 18(2):30, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano62w].

Anonymous:1962:SN

[Ano62z] Anonymous. Science notes. Bulletin of the Atomic Scientists, 18 (1):48, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:SAA

[Ano62-27] Anonymous. Scientists appraise atmospheric tests. Bulletin of the Atomic Scientists, 18(4):33, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:SIW

[Ano62-28] Anonymous. Shelter interest wanes; Peace Corps aid; research exchange; Norway supports ban; names in the news. Bulletin of the Atomic Scientists, 18(6):47–48, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:SD

[Ano62-29] Anonymous. Space developments. Bulletin of the Atomic Scientists, 18(1):46–47, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1962:SR

[Ano62-30] Anonymous. Sweden’s reactor. Bulletin of the Atomic Scientists, 18(5):48, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Aa

[Ano63a] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(1): 40, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ab

[Ano63b] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(2): 48, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ac

[Ano63c] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(3):40, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ad

[Ano63d] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(4):48, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ae

[Ano63e] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(5):48, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Af

[Ano63f] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(6):40, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ag

[Ano63g] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(7): 40, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ah

[Ano63h] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(8): 48, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Ai

[Ano63i] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(9): 48, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:Aj

[Ano63j] Anonymous. Authors. Bulletin of the Atomic Scientists, 19(10): 53, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:CC

[Ano63k] Anonymous. Conversations with the Committee. Bulletin of the Atomic Scientists, 19(3):15, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:EFS

[Ano63l] Anonymous. Editorial: First step — to where? Bulletin of the Atomic Scientists, 19(8):2–3, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:EST

[Ano63m] Anonymous. Editorial: Since ten years. Bulletin of the Atomic Scientists, 19(5):2, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Editorial in support of J. Robert Oppenheimer, written before the announcement of the awarding to him by the Atomic Energy Commission of the Enrico Fermi Prize.

Anonymous:1963:ETB

[Ano63n] Anonymous. Editorial: The test ban. Bulletin of the Atomic Scientists, 19(7):2, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:EPC

[Ano63o] Anonymous. The Eleventh Pugwash Conference [Dubrovnik, Yugoslavia, 20–25 September 1963]. Bulletin of the Atomic Scientists, 19(9):46–48, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:ICM

[Ano63p] Anonymous. International cooperation: Memorandum on cooperation. Bulletin of the Atomic Scientists, 19(7):26–27, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:JKN

[Ano63q] Anonymous. [John Kennedy, 25 November 1963]. Bulletin of the Atomic Scientists, 19(10):1, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:NRTb

[Ano63r] Anonymous. News & reports: Test ban: Choice between risks. Bulletin of the Atomic Scientists, 19(5):37–44, May 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:NRTc

[Ano63s] Anonymous. News & reports: Test ban: “The first concrete measure”. Bulletin of the Atomic Scientists, 19(8):40, 42, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:NRTd

[Ano63t] Anonymous. News & reports: Test ban: The next round. Bulletin of the Atomic Scientists, 19(9):42–43, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:NRTa

[Ano63u] Anonymous. News & reports: Test bans and the Black Box. Bulletin of the Atomic Scientists, 19(1):34–35, January 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:OCE

[Ano63v] Anonymous. Our common enterprise. Bulletin of the Atomic Scientists, 19(7):24, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:PPW

[Ano63w] Anonymous. The Pauling prize: A welcome honor from Norway. Bulletin of the Atomic Scientists, 19(10):18, December 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Don64].

Anonymous:1963:RS

[Ano63x] Anonymous. Radiation standards. Bulletin of the Atomic Scientists, 19(1):37–40, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:RSP

[Ano63y] Anonymous. Report on space programs: Space Science Board study. Bulletin of the Atomic Scientists, 19(5):18–20, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:SPP

[Ano63z] Anonymous. Soviet patent policy. Bulletin of the Atomic Scientists, 19(10):12, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:M

[Ano63-27] Anonymous. To the Moon. Bulletin of the Atomic Scientists, 19 (5):22–24, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:VRD

[Ano63-28] Anonymous. Verification and response in disarmament agreements. Bulletin of the Atomic Scientists, 19(4):44–48, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1963:WWW

[Ano63-29] Anonymous. Writers writing & the written. Bulletin of the Atomic Scientists, 19(9):13, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:APS

[Ano64a] Anonymous. Agricultural products: The 1962 Soviet report. Bulletin of the Atomic Scientists, 20(1):13, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:Aa

[Ano64b] Anonymous. Authors. Bulletin of the Atomic Scientists, 20(1): 40, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:Ab

[Ano64c] Anonymous. Authors. Bulletin of the Atomic Scientists, 20(2): 40, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:Ag

[Ano64d] Anonymous. Authors. Bulletin of the Atomic Scientists, 20(7): 48, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:Ah

[Ano64e] Anonymous. Authors. Bulletin of the Atomic Scientists, 20(8): 48, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:Ai

[Ano64f] Anonymous. Authors. Bulletin of the Atomic Scientists, 20(9): 40, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:Aj

[Ano64g] Anonymous. Authors. Bulletin of the Atomic Scientists, 20(10): 45, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:BAS

[Ano64h] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XX, 1964. Bulletin of the Atomic Scientists, 20(10):46–48, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:BCA

[Ano64i] Anonymous. The budget cuts and the AEC. Bulletin of the Atomic Scientists, 20(3):39–40, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:EGL

[Ano64j] Anonymous. Editorial: Gresham’s Law in Soviet–American exchange. Bulletin of the Atomic Scientists, 20(6):2–3, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:ESI

[Ano64k] Anonymous. Errata: A strategy for innovation. Bulletin of the Atomic Scientists, 20(6):38, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mic64].

Anonymous:1964:IT

[Ano64l] Anonymous. Industry testifies. Bulletin of the Atomic Scientists, 20(4):23, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:KA

[Ano64m] Anonymous. Khrushchev in Astrakhan. Bulletin of the Atomic Scientists, 20(1):12–14, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:OD

[Ano64n] Anonymous. Opening the door. Bulletin of the Atomic Scientists, 20(4):27, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:PSC

[Ano64o] Anonymous. Pugwash on scientific cooperation. Bulletin of the Atomic Scientists, 20(7):41–46, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:PXO

[Ano64p] Anonymous. Pugwash XII: Official statement of the twelfth Pugwash Conference on Science and World Affairs, held in Udaipur, India, January 27 to February 1, 1964. Bulletin of the Atomic Scientists, 20(6):45–48, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1964:PXX

[Ano64q] Anonymous. Pugwash XIII [XIII International Conference on Science and World Affairs, Karlovy Vary, Czechoslovakia, September 13–19, 1964]. Bulletin of the Atomic Scientists, 20(10):43–45, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:Aa

[Ano65a] Anonymous. Authors. Bulletin of the Atomic Scientists, 21(1): 48, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:Ab

[Ano65b] Anonymous. Authors. Bulletin of the Atomic Scientists, 21(2): 48, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:Ac

[Ano65c] Anonymous. Authors. Bulletin of the Atomic Scientists, 21(3):56, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:BAS

[Ano65d] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXI, 1965. Bulletin of the Atomic Scientists, 21(10):46–48, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:EBA

[Ano65e] Anonymous. Einige Briefe aus Amerika. China. (German) [Some letters from America. China]. Neue Wege, 59(??):??, 1965. ISSN 0170-9410.

Anonymous:1965:FAR

[Ano65f] Anonymous. Fermi Award to Robert Wilson. Bulletin of the Atomic Scientists, 21(2):57, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:SSA

[Ano65g] Anonymous. Scientists speak: The American Chemical Society. Statement by mathematicians. FAS on nuclear weapons control. Bulletin of the Atomic Scientists, 21(2):47, February 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1965:WPF

[Ano65h] Anonymous. What the pending Federal bill would do. Bulletin of the Atomic Scientists, 21(4):18, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:BBA

[Ano66a] Anonymous. Books: Bulletin of the Atomic Scientists: Index to Volume XXII, 1966. Bulletin of the Atomic Scientists, 22(10):46– 48, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:CT

[Ano66b] Anonymous. China today. Bulletin of the Atomic Scientists, 22 (6):2–4, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:CSP

[Ano66c] Anonymous. Citizens speak out on proliferation. Bulletin of the Atomic Scientists, 22(9):37–39, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:DA

[Ano66d] Anonymous. The development of agriculture. Bulletin of the Atomic Scientists, 22(6):27, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:E

[Ano66e] Anonymous. The economy. Bulletin of the Atomic Scientists, 22(6):5, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:ECA

[Ano66f] Anonymous. The Editor comments: The atomic bomb secret — fifteen years later. Bulletin of the Atomic Scientists, 22(10):2–3, 25, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:ENa

[Ano66g] Anonymous. Editor’s note. Bulletin of the Atomic Scientists, 22 (2):34, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:ENb

[Ano66h] Anonymous. Editor’s note. Bulletin of the Atomic Scientists, 22 (3):28, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fra66].

Anonymous:1966:ENc

[Ano66i] Anonymous. Editor’s note. Bulletin of the Atomic Scientists, 22 (4):33, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:EPS

[Ano66j] Anonymous. Erratum: Poets and scientists. Bulletin of the Atomic Scientists, 22(1):38, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Corrects transposition of two paragraphs.

Anonymous:1966:INC

[Ano66k] Anonymous. Impressions of the new China. Bulletin of the Atomic Scientists, 22(6):75, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:SCP

[Ano66l] Anonymous. In summary: The Committee’s proposals for arms control and disarmament. Bulletin of the Atomic Scientists, 22(3):34–35, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:IPF

[Ano66m] Anonymous. Internal politics and foreign policy. Bulletin of the Atomic Scientists, 22(6):56–57, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [TW66].

Anonymous:1966:NTC

[Ano66n] Anonymous. The nonproliferation treaties compared. Bulletin of the Atomic Scientists, 22(1):44–45, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:P

[Ano66o] Anonymous. The population. Bulletin of the Atomic Scientists, 22(6):21, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:PX

[Ano66p] Anonymous. Pugwash XV. Bulletin of the Atomic Scientists, 22 (4):42–47, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:SE

[Ano66q] Anonymous. Science and education. Bulletin of the Atomic Scientists, 22(6):40, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1966:SSC

[Ano66r] Anonymous. Scientists speak out on CB weapons. Bulletin of the Atomic Scientists, 22(9):39–40, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The letter is signed by Felix Bloch (Nobel 1952), Konrad E. Bloch (Nobel 1964), James F. Crow, William Doering, Paul Doty, Freeman J. Dyson, John T. Edsall, Bernard Feld, Irwin C. Gunsalus, Robert Hofstadter (Nobel 1961), Arthur Kornberg (Nobel 1959), Fritz Lipmann (Nobel 1953), Robert B. Livingston, Matthew Meselson, Severo Ochoa (Nobel 1959), Ray D. Owen, Keith R. Porter, Charles Price, Eugene Rabinowitch, E. L. Tatum (Nobel 1958), George Wald, and Paul Dudley White.

Anonymous:1966:URN

[Ano66s] Anonymous. The U.N. resolution on nonintervention: December 21, 1945. Bulletin of the Atomic Scientists, 22(5):22, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:ADC

[Ano67a] Anonymous. The ABM debate continues. Bulletin of the Atomic Scientists, 23(6):5, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:ASF

[Ano67b] Anonymous. Arms sales and foreign policy. Bulletin of the Atomic Scientists, 23(7):44–48, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:BAS

[Ano67c] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXIII, 1967. Bulletin of the Atomic Scientists, 23(10):54–56, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:CSN

[Ano67d] Anonymous. A Chinese statement on nuclear proliferation. Bulletin of the Atomic Scientists, 23(5):53–54, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:CEN

[Ano67e] Anonymous. Comments: Editor’s note once again: Space. Bulletin of the Atomic Scientists, 23(3):30, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ure67, Bor68].

Anonymous:1967:CVA

[Ano67f] Anonymous. Comments: Vietnam appeal. Bulletin of the Atomic Scientists, 23(10):45, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:FFS

[Ano67g] Anonymous. Federal funds for science. Bulletin of the Atomic Scientists, 23(2):33–38, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:ICS

[Ano67h] Anonymous. International control in space: 1967. Bulletin of the Atomic Scientists, 23(3):46–48, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:NCW

[Ano67i] Anonymous. Nawapa: A continental water system. Bulletin of the Atomic Scientists, 23(7):8, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:OLF

[Ano67j] Anonymous. An open letter from French scientists. Bulletin of the Atomic Scientists, 23(1):47–48, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:PI

[Ano67k] Anonymous. Planetary investigations. Bulletin of the Atomic Scientists, 23(5):10, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpted from a 1965 National Academy of Sciences report.

Anonymous:1967:PXS

[Ano67l] Anonymous. Pugwash XVI: The statement. Bulletin of the Atomic Scientists, 23(1):44–47, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:RAA

[Ano67m] Anonymous. A report to the American academic community. Bulletin of the Atomic Scientists, 23(4):40–44, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:SJP

[Ano67n] Anonymous. A statement by Japanese physicists. Bulletin of the Atomic Scientists, 23(1):48, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1967:VSS

[Ano67o] Anonymous. Vietnam statements by scientists abroad. Bulletin of the Atomic Scientists, 23(1):47, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:BR

[Ano68a] Anonymous. Books received. Bulletin of the Atomic Scientists, 24(5):39, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:BAS

[Ano68b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXIV, 1968. Bulletin of the Atomic Scientists, 24(10):46–48, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:BVG

[Ano68c] Anonymous. British veto 300 GeV. Bulletin of the Atomic Scientists, 24(7):31, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:CCA

[Ano68d] Anonymous. Current comments: An addendum on Greek resistance. Bulletin of the Atomic Scientists, 24(9):39–40, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano68j].

Anonymous:1968:CCC

[Ano68e] Anonymous. Current comments: Cern accelerator. Bulletin of the Atomic Scientists, 24(10):34, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:CCS

[Ano68f] Anonymous. Current comments: Soviet memorandum of July 1, 1968. Bulletin of the Atomic Scientists, 24(7):30–31, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:EFA

[Ano68g] Anonymous. The Enrico Fermi Award, 1968 [to John Archibald Wheeler]. Bulletin of the Atomic Scientists, 24(9):40, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:FJS

[Ano68h] Anonymous. French and Japanese scientists on war research. Bulletin of the Atomic Scientists, 24(3):35, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter signed by 433 university scientists in France and Japan appealing to American colleagues to refuse to allow their universities to be used for military and secret research..

Anonymous:1968:NB

[Ano68i] Anonymous. New brew. Bulletin of the Atomic Scientists, 24(4): 11, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Two-paragraph excerpt from The Economist, 7 October 1967.

Anonymous:1968:RMR

[Ano68j] Anonymous. Report on the military rule and the Greek universities. Bulletin of the Atomic Scientists, 24(7):20–25, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Ano68d].

Anonymous:1968:SST

[Ano68k] Anonymous. Space science and tight budgets. Bulletin of the Atomic Scientists, 24(6):6–7, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1968:WCB

[Ano68l] Anonymous. Wissenschafter zur chemischen und biologischen Kriegführung aus “The Bulletin of the Atomic Scientists”, September 1968. (German) [Scientists on chemical and biological warfare from “The Bulletin of the Atomic Scientists”, September 1968]. Neue Wege, 62(??):??, 1968. ISSN 0170-9410.

Anonymous:1969:AMS

[Ano69a] Anonymous. Against the misuse of science — an appeal by M.I.T. scientists. Bulletin of the Atomic Scientists, 25(3):8, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Signed by 46 scientists.

Anonymous:1969:BAS

[Ano69b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXV, 1969. Bulletin of the Atomic Scientists, 25(10):45–48, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:CACa

[Ano69c] Anonymous. China after the Cultural Revolution. Bulletin of the Atomic Scientists, 25(2):2–3, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Anonymous:1969:CACb

[Ano69d] Anonymous. China after the cultural revolution: a selection from the Bulletin of the Atomic Scientists. Random House, New York, NY, USA, 1969. vi + 247 pp. LCCN DS777.55 .C44673.

Anonymous:1969:CRO

[Ano69e] Anonymous. The clock retreats once more. Bulletin of the Atomic Scientists, 25(4):48, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:CSS

[Ano69f] Anonymous. Comment: Science and social controls. Bulletin of the Atomic Scientists, 25(5):21, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:CIF

[Ano69g] Anonymous. Comments: Irradiating the fruit fly. Bulletin of the Atomic Scientists, 25(10):34, 42, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:CRW

[Ano69h] Anonymous. Contaminated rain water. Bulletin of the Atomic Scientists, 25(6):43, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:EC

[Ano69i] Anonymous. The economy [of China]. Bulletin of the Atomic Scientists, 25(2):33, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Anonymous:1969:FPC

[Ano69j] Anonymous. Foreign policy [of China]. Bulletin of the Atomic Scientists, 25(2):52, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Anonymous:1969:ORSb

[Ano69k] Anonymous. Other reactions in the science community: Federation of American Scientists. Bulletin of the Atomic Scientists, 25 (6):47–48, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:ORSa

[Ano69l] Anonymous. Other reactions in the science community: Societies for experimental biology. Bulletin of the Atomic Scientists, 25 (6):47, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:PMM

[Ano69m] Anonymous. Part I: The Moon and man. Bulletin of the Atomic Scientists, 25(7):3, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:PIP

[Ano69n] Anonymous. Part II: The politics of spacefaring. Bulletin of the Atomic Scientists, 25(7):16, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:PIF

[Ano69o] Anonymous. Part III: The future of lunar studies. Bulletin of the Atomic Scientists, 25(7):44, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:PIT

[Ano69p] Anonymous. Part IV: The technological impact. Bulletin of the Atomic Scientists, 25(7):69, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1969:PSC

[Ano69q] Anonymous. The political struggle in China. Bulletin of the Atomic Scientists, 25(2):11, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Anonymous:1969:STC

[Ano69r] Anonymous. Science and technology [in China]. Bulletin of the Atomic Scientists, 25(2):79, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Anonymous:1969:SSC

[Ano69s] Anonymous. The seven stages of the Cultural Revolution. Bulletin of the Atomic Scientists, 25(2):16–17, February 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Anonymous:1970:PCS

[Ano70a] Anonymous. 19th Pugwash Conference on Science and World Affairs: The Continuing Committee Report. Bulletin of the Atomic Scientists, 26(4):21–24, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1970:CPS

[Ano70b] Anonymous. Communications: Professor Sternglass, fallout and infant mortality. Bulletin of the Atomic Scientists, 26(3):46, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e].

Anonymous:1970:CA

[Ano70c] Anonymous. The crossing of Antarctica. Bulletin of the Atomic Scientists, 26(10):80, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Anonymous:1970:RAR

[Ano70d] Anonymous. Reports: Age and response to sonic booms. Bulletin of the Atomic Scientists, 26(5):27–28, May 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the Stanford Research Institute Journal, December 1969.

Anonymous:1970:RED

[Ano70e] Anonymous. Reports: Environmental disaster — acts of nature and man: The Peru earthquake: A special study. Bulletin of the Atomic Scientists, 26(8):17–19, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1970:RT

[Ano70f] Anonymous. Reports: Trends. Bulletin of the Atomic Scientists, 26(4):17, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1970:TIU

[Ano70g] Anonymous. Trends: India; university research; anti-pollution. Bulletin of the Atomic Scientists, 26(8):30, October 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:BAS

[Ano71a] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXVII, 1971. Bulletin of the Atomic Scientists, 27(10):57–60, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:DAL

[Ano71b] Anonymous. Disarmament and arms limitation: Pugwash assessment. Bulletin of the Atomic Scientists, 27(3):20–23, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:ECP

[Ano71c] Anonymous. Energy crisis: Part II. Bulletin of the Atomic Scientists, 27(9):37, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:FFR

[Ano71d] Anonymous. Fission and fusion reactors: The Alfven memorandum. Bulletin of the Atomic Scientists, 27(7):36, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:SSS

[Ano71e] Anonymous. Science and society symposium: Prospects for survival. Bulletin of the Atomic Scientists, 27(5):19–22, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:SRC

[Ano71f] Anonymous. A special report: On the control of science: Four views. Bulletin of the Atomic Scientists, 27(10):23, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:SAE

[Ano71g] Anonymous. Summary: AEC environmental statement, Radioactive Waste Repository, Lyons, Kansas. Bulletin of the Atomic Scientists, 27(6):33–34, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:TPA

[Ano71h] Anonymous. Trends: Patent applications. Bulletin of the Atomic Scientists, 27(1):33, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:RC

[Ano72a] Anonymous. 16th Rochester Conference. Bulletin of the Atomic Scientists, 28(7):57–58, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:BDC

[Ano72b] Anonymous. The Biological Disarmament Convention. Bulletin of the Atomic Scientists, 28(4):11, 17, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:BR

[Ano72c] Anonymous. Books received. Bulletin of the Atomic Scientists, 28(5):41, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:BAS

[Ano72d] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXVIII, 1972. Bulletin of the Atomic Scientists, 28(10):53–56, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:EAA

[Ano72e] Anonymous. End of Apollo: The ambiguous epic. Bulletin of the Atomic Scientists, 28(10):39–44, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:RHA

[Ano72f] Anonymous. Rabinowitch honored [by the American Academy of Arts and Sciences]. Bulletin of the Atomic Scientists, 28(6):40, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1972:SVC

[Ano72g] Anonymous. Scientist vs. citizen: The Shoreham controversy. Bulletin of the Atomic Scientists, 28(6):24, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fes73].

Anonymous:1972:SRW

[Ano72h] Anonymous. A special report: What happened at Stockholm: IV. Declaration on the human environment: A statement of principles. Bulletin of the Atomic Scientists, 28(7):33–34, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment about omitted paragraph in point 19 [Red74].

Anonymous:1972:WRO

[Ano72i] Anonymous. We reset our clock. Bulletin of the Atomic Scientists, 28(6):3, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:BAS

[Ano73a] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXIX, 1973. Bulletin of the Atomic Scientists, 29(10):61–64, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:CEL

[Ano73b] Anonymous. Communications: Evgeny Levich — an appeal. Bulletin of the Atomic Scientists, 29(10):2, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:ER

[Ano73c] Anonymous. Eugene Rabinowitch 1901–1973. Bulletin of the Atomic Scientists, 29(6):2–3, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:MER

[Ano73d] Anonymous. In memoriam: Eugene Rabinowitch 1901–1973: A voice of conscience for the Atomic Age is stilled. Bulletin of the Atomic Scientists, 29(6):c1, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:NM

[Ano73e] Anonymous. The new Moon. Bulletin of the Atomic Scientists, 29 (9):4, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:NMP

[Ano73f] Anonymous. The new Moon, Part II. Bulletin of the Atomic Scientists, 29(10):18, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:NCM

[Ano73g] Anonymous. A note on carbon monoxide. Bulletin of the Atomic Scientists, 29(5):23, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1973:SEC

[Ano73h] Anonymous. Summary of EPA clean air proposals for Los Angeles area. Bulletin of the Atomic Scientists, 29(4):7, April 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:BAS

[Ano74a] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXX, 1974. Bulletin of the Atomic Scientists, 30(10):45–48, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:CRF

[Ano74b] Anonymous. Communications: In re “A ‘Free’ Slave”. Bulletin of the Atomic Scientists, 30(2):45, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jan73]. The letter is signed by 34 faculty members of the Department of Radiation Biology and Biophysics at the University of Rochester.

Anonymous:1974:CRO

[Ano74c] Anonymous. Correction: [the race to oblivion]. Bulletin of the Atomic Scientists, 30(9):48, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei74].

Anonymous:1974:CMC

[Ano74d] Anonymous. Critical mass on Capitol Hill. Bulletin of the Atomic Scientists, 30(4):8–9, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:ERA

[Ano74e] Anonymous. Errata: “The Role of Anti-Matter in Big-Bang Cosmology”. Bulletin of the Atomic Scientists, 30(2):45, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Repairs an incorrect reference in [Ste74] to Herman Friedman to the correct Alexandre Friedmann.

Anonymous:1974:HKJ

[Ano74f] Anonymous. Harry Kalven, Jr. 1914–1974. Bulletin of the Atomic Scientists, 30(10):5, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:MM

[Ano74g] Anonymous. Minutes to midnight. Bulletin of the Atomic Scientists, 30(5):6–9, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:NMP

[Ano74h] Anonymous. The new Moon — Part III. Bulletin of the Atomic Scientists, 30(3):30, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:PRC

[Ano74i] Anonymous. Pugwash 1973: Report of the Continuing Committee: Despite detente, a mounting danger to world peace, health and security. Bulletin of the Atomic Scientists, 30(4):24–31, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1974:SN

[Ano74j] Anonymous. Staff note. Bulletin of the Atomic Scientists, 30 (3):3, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Note on the resignation of editor Richard S. Lewis and his replacement by Samuel H. Day, Jr..

Anonymous:1975:BFN

[Ano75a] Anonymous. Bernard T. Feld named Editor-in-Chief. Bulletin of the Atomic Scientists, 31(9):7, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1975:BAS

[Ano75b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXI, 1975. Bulletin of the Atomic Scientists, 31(10):60–63, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1975:CDC

[Ano75c] Anonymous. The Cocoyoc Declaration: A call for reform of the international economic order. Bulletin of the Atomic Scientists, 31(3):6–10, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1975:CNT

[Ano75d] Anonymous. Correction: [number of Trident submarines]. Bulletin of the Atomic Scientists, 31(6):4, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ros75].

Anonymous:1975:NTP

[Ano75e] Anonymous. A note of thanks [to publisher James P. Cahill]. Bulletin of the Atomic Scientists, 31(7):6, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1975:NB

[Ano75f] Anonymous. Nothing but the best. Bulletin of the Atomic Scientists, 31(2):15, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1975:NAL

[Ano75g] Anonymous. Nukleare Abrüstung laut “Bulletin of the Atomic Scientists” ein Trugbild. (German) [Nuclear disarmament according to “Bulletin of the Atomic Scientists” a mirage]. Neue Wege, ??(??):??, 1975. ISSN 0170-9410.

Anonymous:1975:SL

[Ano75h] Anonymous. The silver lining. Bulletin of the Atomic Scientists, 31(2):31, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Brief comment on the poor absorption of plutonium by plants and fish.

Anonymous:1975:WG

[Ano75i] Anonymous. A weapons glossary. Bulletin of the Atomic Scientists, 31(4):13, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1976:BAS

[Ano76a] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXII, 1976. Bulletin of the Atomic Scientists, 32(10):60–63, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1976:DRW

[Ano76b] Anonymous. Development, resources and world security. Bulletin of the Atomic Scientists, 32(3):24–29, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). 1976 Pugwash Conference report.

Anonymous:1976:PSS

[Ano76c] Anonymous. Peace science society. Bulletin of the Atomic Scientists, 32(7):6, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dum76].

Anonymous:1976:POD

[Ano76d] Anonymous. Public opinion and disarmament: a European initiative. Bulletin of the Atomic Scientists, 32(10):41, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABa

[Ano77a] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(2):61, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABb

[Ano77b] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(3):70, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABc

[Ano77c] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(4):61, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABd

[Ano77d] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(5):60, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABe

[Ano77e] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(6):57, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABf

[Ano77f] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(7):62, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABg

[Ano77g] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(8):62, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABh

[Ano77h] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(9):48, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:YABi

[Ano77i] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 33(10):57, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:APA

[Ano77j] Anonymous. An action program against cancer. Bulletin of the Atomic Scientists, 33(3):30, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:BAS

[Ano77k] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXIII, 1977. Bulletin of the Atomic Scientists, 33(10):61–64, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:CER

[Ano77l] Anonymous. The Cambridge Experimentation Review Board: How a citizens group helped a city council set safety standards for genetic research. Bulletin of the Atomic Scientists, 33(5): 22–27, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1977:USS

[Ano77m] Anonymous. U.S.–Soviet SALT proposals. Bulletin of the Atomic Scientists, 33(5):9, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Short excerpts from the New York Times, April 1, 3, 1977.

Anonymous:1977:UAN

[Ano77n] Anonymous. U.S. Army’s new radiation casualty criteria. Bulletin of the Atomic Scientists, 33(7):35, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:YABa

[Ano78a] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 34(1):52, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:YABb

[Ano78b] Anonymous. 10, 20, 30 years ago in the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 34(2):60, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:A

[Ano78c] Anonymous. Apologies. Bulletin of the Atomic Scientists, 34(9): 59, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [dV78].

Anonymous:1978:BRBb

[Ano78d] Anonymous. Book reviews: Overcoming Legal Uncertainties About the Use of Solar Energy Systems, by William A. Thomas, Alan S. Miller, and Richard L. Robbins. Bulletin of the Atomic Scientists, 34(8):49, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:BRBa

[Ano78e] Anonymous. Book reviews: Soviet Science, by Zhores A. Medvedev. Bulletin of the Atomic Scientists, 34(8):48, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:BRBc

[Ano78f] Anonymous. Book reviews: The Effects of Nuclear Weapons, edited by Samuel Glasstone and Phillip J. Dolan. Bulletin of the Atomic Scientists, 34(8):49, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:BBE

[Ano78g] Anonymous. Books: The Environmental Protection Agency Needs Congressional Guidance and Support to Guard the Public in a Period of Radiation Proliferation. Bulletin of the Atomic Scientists, 34(7):50, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:BAS

[Ano78h] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXIV, 1978. Bulletin of the Atomic Scientists, 34(10):59–64, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:CSP

[Ano78i] Anonymous. Commentary: Scientists protest trials in USSR. Bulletin of the Atomic Scientists, 34(7):52–54, September 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:C

[Ano78j] Anonymous. Corrections. Bulletin of the Atomic Scientists, 34 (10):58, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hud78, Mor78].

Anonymous:1978:RTBa

[Ano78k] Anonymous. The reading table: Cancer Crusade: The Story of the National Cancer Act of 1971, by Richard A. Rettig; An Anatomy of Risk, by William D. Rowe; The Picture of Health: Environmental Sources of Disease, by Erik P. Eckholm. Bulletin of the Atomic Scientists, 34(2):63–64, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:RTBc

[Ano78l] Anonymous. The reading table: Pugwash on Self-Reliance, edited by W. K. Chagula, B. T. Feld, A. Parthasarathi, P. J. Lavakare; Perspectives in the Sociology of Science, edited by Stuart S. Blume; Disarmament Times. Bulletin of the Atomic Scientists, 34(4):63, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:RTBe

[Ano78m] Anonymous. The reading table: Rays of Hope: The Transition to a Post-Petroleum World, by Denis Hayes; Opportunities for Disarmament: A Preview of the 1978 United Nations Special Session on Disarmament, edited by Jane M. O. Sharp; “What is the World Made of? Atoms, Leptons, Quarks and Other Tantalizing Particles”, by Gerald Feinberg; A Constitution for the Federation of Earth. Bulletin of the Atomic Scientists, 34(6):63, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:RTBb

[Ano78n] Anonymous. The reading table: Soviet Sociology of Science, by Linda L. Lubrano; Weapons of Mass Destruction and the Environment; Managing the Flow of Technology, by Thomas J. Allen; The Pocket Calculator Game Book, by Edwin Schlossbert and John Brockman; The Imposter, by Helen McCloy; Radioactive Wastes: Management and Regulation, by Mason Willrich and Richard K. Lester; Transporting Natural Gas from the Artic: The Alternative Systems, by Walter J. Mead with George W. Rogers and Rufus Z. Smith; The Direct Connection, by John R. Hadd; Federal Funding of Civilian Research and Development, by Michael Michaelis; The United Nations Disarmament Yearbook; YV 88: An Eco-Fiction of Tomorrow, by Christopher Swan and Chet Roaman; Business and Environment: Toward Common Ground, edited by H. Jeffrey Leonard, J. Clarence Davies III, and Gordon Binder; Toward a Human World Order, by Gerald Mische and Patricia Mische; Until the Sun Dies, by Robert Jastrow; The Inner Planets, by Clark R. Chapman; Decision-Making in the Environmental Protection Agency. Bulletin of the Atomic Scientists, 34(3):63–64, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:RTBd

[Ano78o] Anonymous. The reading table: The Samizdat Register, by Roy A. Medvedev; U.S. Energy Conservation Could Benefit From Experiences of Other Countries, Prosperity Without Guns: The Economic Impact of Reductions in Defense Spending, by B. G. Lall; Essay of an Information Scientist, Volumes 1 and 2, by Eugene Garfield. Bulletin of the Atomic Scientists, 34(5):55, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:RTF

[Ano78p] Anonymous. The reading table: Free environment for science: Scholarly Freedom and Human Rights; Crossed wires: The Social Impact of the Telephone, by Ithiel de Sola Pool; Social systems and biology: Biology as a Social Weapon. Bulletin of the Atomic Scientists, 34(1):55, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1978:YFO

[Ano78q] Anonymous. Yuri F. Orlov: sentenced on May 18, 1978. Bulletin of the Atomic Scientists, 34(7):6, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:BRB

[Ano79a] Anonymous. Book reviews: Space — Battlefield of the Future?, by Bhupendra M. Jasani. Bulletin of the Atomic Scientists, 35(2): 49, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:BAS

[Ano79b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXV, 1979. Bulletin of the Atomic Scientists, 35(10):59–64, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:CLW

[Ano79c] Anonymous. China looks West. Bulletin of the Atomic Scientists, 35(1):55, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint of editorial in Nature.

Anonymous:1979:CMM

[Ano79d] Anonymous. Commentary from Mexico: 1979 message from Pugwash. Bulletin of the Atomic Scientists, 35(9):62–63, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:ENC

[Ano79e] Anonymous. Editor’s note: [Comprehensive Nuclear Test Ban]. Bulletin of the Atomic Scientists, 35(2):60, February 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:ELSa

[Ano79f] Anonymous. Excerpts: Leo Szilard: His version of the facts. Bulletin of the Atomic Scientists, 35(2):37–40, February 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:ELSb

[Ano79g] Anonymous. Excerpts: Leo Szilard: His version of the facts: Part II. Bulletin of the Atomic Scientists, 35(3):55–59, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:GA

[Ano79h] Anonymous. Geneva appeal. Bulletin of the Atomic Scientists, 35(6):12, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:LSH

[Ano79i] Anonymous. Leo Szilard: His version of the facts. Bulletin of the Atomic Scientists, 35(4):28–32, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:MIU

[Ano79j] Anonymous. Montana initiative [on use of nuclear power]. Bulletin of the Atomic Scientists, 35(1):7, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:NGN

[Ano79k] Anonymous. No general need apply. Bulletin of the Atomic Scientists, 35(2):4, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:ONS

[Ano79l] Anonymous. Old nuclear sites. Bulletin of the Atomic Scientists, 35(5):48, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:PFL

[Ano79m] Anonymous. Part four: Leo Szilard: His version of the facts. Bulletin of the Atomic Scientists, 35(5):34–35, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:USP

[Ano79n] Anonymous. U.S. Studnet Pugwash Conference. Bulletin of the Atomic Scientists, 35(5):50, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1979:WPA

[Ano79o] Anonymous. Wanted: Peace abstractors. Bulletin of the Atomic Scientists, 35(3):87, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:AM

[Ano80a] Anonymous. April: Moscow. Bulletin of the Atomic Scientists, 36(4):65, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:ACM

[Ano80b] Anonymous. August: Cambridge, Mass. Bulletin of the Atomic Scientists, 36(5):65, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:AVa

[Ano80c] Anonymous. August: Venice. Bulletin of the Atomic Scientists, 36(4):65, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:AVb

[Ano80d] Anonymous. August: Venice. Bulletin of the Atomic Scientists, 36(5):65, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:BAS

[Ano80e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXVI, 1980. Bulletin of the Atomic Scientists, 36(10):60–64, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:CSC

[Ano80f] Anonymous. Commentary: From Science Council of Japan. Bulletin of the Atomic Scientists, 36(8):59, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:CNP

[Ano80g] Anonymous. Commentary from The Netherlands: Pugwash 1980. Bulletin of the Atomic Scientists, 36(9):9, November 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:CDV

[Ano80h] Anonymous. Correction: Dioxin and the Vietnam veteran. Bulletin of the Atomic Scientists, 36(1):64, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DD79].

Anonymous:1980:EPI

[Ano80i] Anonymous. Erratum: [Pershing II: the Army’s strategic weapon]. Bulletin of the Atomic Scientists, 36(10):58, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pai80a].

Anonymous:1980:ERP

[Ano80j] Anonymous. Erratum: [Research and politics in the People’s Republic of China]. Bulletin of the Atomic Scientists, 36(7):59, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zho79].

Anonymous:1980:EBP

[Ano80k] Anonymous. Excerpts from Report of the President’s Commission on the Accident at Three Mile Island: The Need for Change: The Legacy of TMI. Bulletin of the Atomic Scientists, 36(1):24–31, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:JWP

[Ano80l] Anonymous. Josephine Wertheim Pomerance, 1911–1980. Bulletin of the Atomic Scientists, 36(8):7, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:LEU

[Ano80m] Anonymous. Letter to the Editor: On U.S.–Soviet exchange programs. Bulletin of the Atomic Scientists, 36(8):59–60, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:MV

[Ano80n] Anonymous. May: Vancouver. Bulletin of the Atomic Scientists, 36(4):65, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:MAB

[Ano80o] Anonymous. Memorial appeal: [bust of Bertrand Russell in gardens of Red Lion Square, London]. Bulletin of the Atomic Scientists, 36(5):64, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:SS

[Ano80p] Anonymous. MIT summer session. Bulletin of the Atomic Scientists, 36(3):63, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:SRP

[Ano80q] Anonymous. Special report: Part one: Nuclear power in 1980. Bulletin of the Atomic Scientists, 36(1):17, January 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1980:TD

[Ano80r] Anonymous. That Trident decision. Bulletin of the Atomic Scientists, 36(9):30, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from The Guardian, 27 July 1980.

Anonymous:1981:ANY

[Ano81a] Anonymous. April: New York. Bulletin of the Atomic Scientists, 37(3):64, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:AVC

[Ano81b] Anonymous. August 1981: Vail, Colorado. Bulletin of the Atomic Scientists, 37(3):64, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:BRB

[Ano81c] Anonymous. Book review: Fear of Science — Trust in Science: Conditions for Change in the Climate of Opinion, edited by A. S. Markovits and K. W. Deutsch. Bulletin of the Atomic Scientists, 37(8):40–41, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:BAS

[Ano81d] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXVII, 1981. Bulletin of the Atomic Scientists, 37(10):60–64, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:EU

[Ano81e] Anonymous. Errata: “It’s Up to Us”. Bulletin of the Atomic Scientists, 37(6):65, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hod81a].

Anonymous:1981:FRB

[Ano81f] Anonymous. Film reviews: Peace in Search of Makers, produced by Jay Bender; The National Nuclear Debate; Models and Methods: A Study in Appropriate Technology, produced by Donald R. Ham. Bulletin of the Atomic Scientists, 37(3):48–49, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:J

[Ano81g] Anonymous. June: MIT. Bulletin of the Atomic Scientists, 37 (3):64, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:JNH

[Ano81h] Anonymous. June: New Haven. Bulletin of the Atomic Scientists, 37(4):65, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:LVD

[Ano81i] Anonymous. London: Vigil for disarmament. Bulletin of the Atomic Scientists, 37(7):64, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:PED

[Ano81j] Anonymous. Pugwash on Europe’s dilemma. Bulletin of the Atomic Scientists, 37(2):5, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:SNA

[Ano81k] Anonymous. Shadows of the nuclear age. Bulletin of the Atomic Scientists, 37(4):50, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:SAA

[Ano81l] Anonymous. Summer: Ann Arbor. Bulletin of the Atomic Scientists, 37(3):64, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1981:VER

[Ano81m] Anonymous. The voice of Eugene Rabinowitch. Bulletin of the Atomic Scientists, 37(1):14, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:A

[Ano82a] Anonymous. Addendum. Bulletin of the Atomic Scientists, 38 (5):62, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mog81].

Anonymous:1982:BAS

[Ano82b] Anonymous. Bulletin of the Atomic Scientists: Index to Volume XXXVIII, 1982. Bulletin of the Atomic Scientists, 38(10):76– 80, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:BAC

[Ano82c] Anonymous. Bulletin: Arms control advocates protest harassment of Soviet peace group. Bulletin of the Atomic Scientists, 38 (9):62–63, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:BJB

[Ano82d] Anonymous. Bulletin: “Japan’s Biological Weapons: 1930–1945” — an update. Bulletin of the Atomic Scientists, 38(8):62, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GPR81].

Anonymous:1982:CFE

[Ano82e] Anonymous. Conference on the Fate of the Earth. Bulletin of the Atomic Scientists, 38(8):63–64, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:ESR

[Ano82f] Anonymous. Environmental and safety risk management. Bulletin of the Atomic Scientists, 38(4):18, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:Ea

[Ano82g] Anonymous. Errata. Bulletin of the Atomic Scientists, 38(7): 63, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei82a, Fri82].

Anonymous:1982:Eb

[Ano82h] Anonymous. Erratum. Bulletin of the Atomic Scientists, 38(9): 57, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zac82].

Anonymous:1982:EAS

[Ano82i] Anonymous. Erratum: [Asian Security in the 1980s]. Bulletin of the Atomic Scientists, 38(5):54, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dre82b].

Anonymous:1982:EDN

[Ano82j] Anonymous. Erratum: “Demise of nuclear deterrence?”. Bulletin of the Atomic Scientists, 38(1):42, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sno81].

Anonymous:1982:ESB

[Ano82k] Anonymous. Erratum: “Science in the Brezhnev era. Bulletin of the Atomic Scientists, 38(3):44, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra82a].

Anonymous:1982:ENG

[Ano82l] Anonymous. Erratum: “The name of the game: Half-SAFE”. Bulletin of the Atomic Scientists, 38(6):40, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DM82].

Anonymous:1982:ENW

[Ano82m] Anonymous. Ethics and nuclear weapons: a play. Bulletin of the Atomic Scientists, 38(3):42, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:FB

[Ano82n] Anonymous. The fate of the book. Bulletin of the Atomic Scientists, 38(8):63, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Comment on Jonathan Schnell’s book The Fate of the Earth.

Anonymous:1982:GZP

[Ano82o] Anonymous. Ground Zero/Phase 2. Bulletin of the Atomic Scientists, 38(7):64–65, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:PCR

[Ano82p] Anonymous. Palme Commission report. Bulletin of the Atomic Scientists, 38(7):65, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:PM

[Ano82q] Anonymous. The problems of measurement. Bulletin of the Atomic Scientists, 38(7):23–24, August/September 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:REW

[Ano82r] Anonymous. Rabinowitch Essay winners. Bulletin of the Atomic Scientists, 38(2):2, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:SSF

[Ano82s] Anonymous. Sakharov scholarship fund. Bulletin of the Atomic Scientists, 38(6):72, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1982:WWN

[Ano82t] Anonymous. Warsaw: We now appeal. Bulletin of the Atomic Scientists, 38(9):3–4, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:AAS

[Ano83a] Anonymous. Addenda: Anthrax in Sverdlovsk? and Afghanistan. Bulletin of the Atomic Scientists, 39(7):64, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [A N83].

Anonymous:1983:ABK

[Ano83b] Anonymous. Addenda: [In the King of Prussia, You can’t have it both ways, and Security and scientific communication]. Bulletin of the Atomic Scientists, 39(8):56, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dow83f, Ung83, Cha83a].

Anonymous:1983:ANP

[Ano83c] Anonymous. Addendum: [Nuclear power and nuclear weapons: the connection is dangerous]. Bulletin of the Atomic Scientists, 39(3):64, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol83b].

Anonymous:1983:AIR

[Ano83d] Anonymous. Amnesty International Report 1982. Bulletin of the Atomic Scientists, 39(1):54–55, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:BAS

[Ano83e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 39, 1983. Bulletin of the Atomic Scientists, 39(10):61–65, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:BTR

[Ano83f] Anonymous. Bulletin: A talk with Randall Kehler. Bulletin of the Atomic Scientists, 39(1):53–54, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:BNW

[Ano83g] Anonymous. Bulletin: Nuclear Weapons Freeze Campaign. Bulletin of the Atomic Scientists, 39(2):57, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:BNC

[Ano83h] Anonymous. Bulletin: The November 11 convocation. Bulletin of the Atomic Scientists, 39(2):57, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:CN

[Ano83i] Anonymous. Course on nonviolence. Bulletin of the Atomic Scientists, 39(2):41, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:E

[Ano83j] Anonymous. Erratum. Bulletin of the Atomic Scientists, 39(2): 41, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Agn82b].

Anonymous:1983:EBF

[Ano83k] Anonymous. Erratum: No First Use. Bulletin of the Atomic Scientists, 39(10):44, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dow83d].

Anonymous:1983:EEI

[Ano83l] Anonymous. Erratum: ‘Electricity and industry’. Bulletin of the Atomic Scientists, 39(5):62, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [MA83].

Anonymous:1983:EIM

[Ano83m] Anonymous. Erratum: Inside the mushroom cloud. Bulletin of the Atomic Scientists, 39(5):58, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tsi83i].

Anonymous:1983:ESS

[Ano83n] Anonymous. Euromissiles: special section. Bulletin of the Atomic Scientists, 39(10):11, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:FAB

[Ano83o] Anonymous. Forum award to Bulletin. Bulletin of the Atomic Scientists, 39(3):3, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:PNL

[Ano83p] Anonymous. Perspective: A new look. Bulletin of the Atomic Scientists, 39(7):2, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:SRF

[Ano83q] Anonymous. STS research fellowships. Bulletin of the Atomic Scientists, 39(10):47, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1983:SC

[Ano83r] Anonymous. Summer course. Bulletin of the Atomic Scientists, 39(3):64, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:AFB

[Ano84a] Anonymous. Addendum: [film U.S. versus U.S.S.R: Who’s Ahead]. Bulletin of the Atomic Scientists, 40(5):59, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dow84a].

Anonymous:1984:AK

[Ano84b] Anonymous. Alfred Kastler, 1902–1984. Bulletin of the Atomic Scientists, 40(3):18, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:BAS

[Ano84c] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 40, 1984. Bulletin of the Atomic Scientists, 40(10):36–40, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:CRN

[Ano84d] Anonymous. Conflict resolution: [1984 National Conference on Peacemaking and Conflict Resolution, University of Missouri–St. Louis, 20–23 September 1984]. Bulletin of the Atomic Scientists, 40(7):45, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:ESE

[Ano84e] Anonymous. Editorial: Sakharov and East–West relations. Bulletin of the Atomic Scientists, 40(7):2, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [vL85].

Anonymous:1984:EB

[Ano84f] Anonymous. Englebert Broda, 1910–1983. Bulletin of the Atomic Scientists, 40(3):18, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:EBW

[Ano84g] Anonymous. Erratum: [Nuclear Weapons Databook: Volume 1: U.S. Nuclear Forces and Capabilities]. Bulletin of the Atomic Scientists, 40(10):31, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cal84a].

Anonymous:1984:EBS

[Ano84h] Anonymous. Erratum: [Strategic Studies and Public Policy: The American Experience]. Bulletin of the Atomic Scientists, 40(8): 42, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Roc83a].

Anonymous:1984:SUD

[Ano84i] Anonymous. From the Soviet Union — a different view. Bulletin of the Atomic Scientists, 40(8):31, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:ICA

[Ano84j] Anonymous. Increasing concern [about tritium in the workplace]. Bulletin of the Atomic Scientists, 40(3):57, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:MAC

[Ano84k] Anonymous. Meeting announcement and call for papers: [International Association for Impact Assessment, New York City, 24–25 May 1984]. Bulletin of the Atomic Scientists, 40(1):41, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:JDP

[Ano84l] Anonymous. Not just for Democrats — a peace strategy. Bulletin of the Atomic Scientists, 40(6):9–10, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:RG

[Ano84m] Anonymous. Resource guide. Bulletin of the Atomic Scientists, 40(10):27s–32s [various between 20 and 21], December 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:RRC

[Ano84n] Anonymous. A risk reduction center. Bulletin of the Atomic Scientists, 40(6):28–29, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:RY

[Ano84o] Anonymous. Ruth Young. Bulletin of the Atomic Scientists, 40 (7):2, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1984:UCP

[Ano84p] Anonymous. U.S.-controlled plutonium in Superphénix. Bulletin of the Atomic Scientists, 40(9):33, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:AAA

[Ano85a] Anonymous. American Academy of Arts and Sciences: George Kistiakowsky Visiting Scholar Program. Bulletin of the Atomic Scientists, 41(4):43, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:BNBb

[Ano85b] Anonymous. Book note: Nuclear Deterrence: Ethics and Strategy, edited by Russell Hardin, John J. Mearsheimer, Gerald Dworkin, and Robert Goodin. Bulletin of the Atomic Scientists, 41(10):52, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:BNBc

[Ano85c] Anonymous. Book notes: Nuclear Strategy, Arms Control, and the Future, edited by P. Edward Haley, David M. Keithly, and Jack Merritt; Thinking about Nuclear Weapons: Analyses and Prescriptions, edited by Fred Holroyd. Bulletin of the Atomic Scientists, 41(11):62, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:BNBa

[Ano85d] Anonymous. Book notes: Search for Sanity: The Politics of Nuclear Weapons and Disarmament, edited by Paul Joseph and Simon Rosenblum; A Guide to Nuclear Power Technology by Frank J. Rahn, Achilles G. Adamantiades, John E. Kenton, and Chaim Braun; The Omni Interviews, edited by Pamela Weintraub; Citizen Participation in Science Policy, edited by James C. Peterson. Bulletin of the Atomic Scientists, 41(5):50–51, May 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:BAS

[Ano85e] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 41, 1985. Bulletin of the Atomic Scientists, 41(11):68–72, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:CS

[Ano85f] Anonymous. Conference statement. Bulletin of the Atomic Scientists, 41(4):50, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:ESP

[Ano85g] Anonymous. Excerpts from the 1985 statement of the Pugwash Council. Bulletin of the Atomic Scientists, 41(10):45, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:FAR

[Ano85h] Anonymous. Fermi award to Robert Wilson [and Georges Vendreyès]. Bulletin of the Atomic Scientists, 41(2):57, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:SPS

[Ano85i] Anonymous. For the species and the planet: a statement in support of the five continent peace initiative. Bulletin of the Atomic Scientists, 41(10):4–5, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:ETS

[Ano85j] Anonymous. From the Editors: A tentative step forward. Bulletin of the Atomic Scientists, 41(3):2, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:EEG

[Ano85k] Anonymous. From the Editors: Enter Gorbachev. Bulletin of the Atomic Scientists, 41(5):2, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:EHB

[Ano85l] Anonymous. From the Editors: Harrison Brown named Editorin-Chief. Bulletin of the Atomic Scientists, 41(2):2, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:ENP

[Ano85m] Anonymous. From the Editors: Nobel for physicians. Bulletin of the Atomic Scientists, 41(11):2, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cal86b].

Anonymous:1985:ENA

[Ano85n] Anonymous. From the Editors: Nuclear allergies. Bulletin of the Atomic Scientists, 41(6):2, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:EOS

[Ano85o] Anonymous. From the Editors: Official secrecy fostered coverup. Bulletin of the Atomic Scientists, 41(4):2, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:ERP

[Ano85p] Anonymous. From the Editors: Risk peace. Bulletin of the Atomic Scientists, 41(1):2, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:EST

[Ano85q] Anonymous. From the Editors: Stop testing. Bulletin of the Atomic Scientists, 41(9):2, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:EES

[Ano85r] Anonymous. From the Editors: The eleventh summit. Bulletin of the Atomic Scientists, 41(10):2, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:ETA

[Ano85s] Anonymous. From the Editors: Two aspects of security. Bulletin of the Atomic Scientists, 41(8):2, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:GAC

[Ano85t] Anonymous. General Advisory Committee Report excerpts. Bulletin of the Atomic Scientists, 41(1):32–34, January 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:HBR

[Ano85u] Anonymous. Hans A. Bethe receives Vannevar Bush Award. Bulletin of the Atomic Scientists, 41(6):55, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:HSJ

[Ano85v] Anonymous. Herbert Scoville, Jr., 1915–1985. Bulletin of the Atomic Scientists, 41(9):57, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:MGA

[Ano85w] Anonymous. Members of the General Advisory Committee. Bulletin of the Atomic Scientists, 41(1):34, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:PWN

[Ano85x] Anonymous. Pugwash workshop on nuclear forces. Bulletin of the Atomic Scientists, 41(3):53, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:WSJ

[Ano85y] Anonymous. Western scientists join Soviet comet mission. Bulletin of the Atomic Scientists, 41(3):19, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:AM

[Ano86a] Anonymous. Alva Myrdal, 1902–1986. Bulletin of the Atomic Scientists, 42(4):56, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:Aa

[Ano86b] Anonymous. Announcements. Bulletin of the Atomic Scientists, 42(8):56, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:Ab

[Ano86c] Anonymous. Announcements. Bulletin of the Atomic Scientists, 42(9):57, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ASP

[Ano86d] Anonymous. Anti-SDI pledge. Bulletin of the Atomic Scientists, 42(1):28, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ACT

[Ano86e] Anonymous. Appendix A: a comparison of the two reports. Bulletin of the Atomic Scientists, 42(6):25–27, June/July 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BLE

[Ano86f] Anonymous. Behind the lobbying effort for weaker sanctions. Bulletin of the Atomic Scientists, 42(3):11, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BSN

[Ano86g] Anonymous. Blocking the spread of nuclear weapons. Bulletin of the Atomic Scientists, 42(6):23, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BNBb

[Ano86h] Anonymous. Book note: Ending Hunger: An Idea Whose Time Has Come, by The Hunger Project. Bulletin of the Atomic Scientists, 42(2):46, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BNBe

[Ano86i] Anonymous. Book notes: Peace Resource Book 1986; Arms Production in the Third World, edited by Michael Brzoska and Thomas Ohlson; Weapons in Space, edited by Franklin A. Long, Donald Hafner, and Jeffrey Boutwell; Empty Promise: The Growing Case against Star Wars, edited by John Tirman; Peace and War: Cross-Cultural Perspectives, edited by Mary LeCron Foster and Robert A. Rubinstein; Planet Earth in Jeopardy, by Lydia Dotto. Bulletin of the Atomic Scientists, 42(10):49–50, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BNBd

[Ano86j] Anonymous. Book notes: The Heavy Dancers, by E. P. Thompson; The 1987 Defense Budget, by Joshua M. Epstein. Bulletin of the Atomic Scientists, 42(9):54, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BNBa

[Ano86k] Anonymous. Book notes: Verification and Arms Control, edited by William C. Potter; The Verification Challenge: Problems and Promise of Strategic Nuclear Arms Control Verification, by Richard Scribner, William Metz, and Theodore Ralston. Bulletin of the Atomic Scientists, 42(1):52, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BNBc

[Ano86l] Anonymous. Book notes: World Armaments and Disarmament: SIPRI Yearbook 1986; Lost at the Frontier: U.S. Science and Technology Policy Adrift, by Deborah Shapley and Rustum Roy. Bulletin of the Atomic Scientists, 42(8):53, October 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BAS

[Ano86m] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 42, 1986. Bulletin of the Atomic Scientists, 42(10):53–56, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:BWA

[Ano86n] Anonymous. Bulletin wins award. Bulletin of the Atomic Scientists, 42(5):48, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:CMS

[Ano86o] Anonymous. A casual mention of Sweden’s “atomic armament” in 1960. Bulletin of the Atomic Scientists, 42(3):33, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Ny Teknik, 2 May 1985, page 8.

Anonymous:1986:CDD

[Ano86p] Anonymous. Community dependence on defense spending: a poll. Bulletin of the Atomic Scientists, 42(8):31, October 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:CCC

[Ano86q] Anonymous. Computer companies comment on sales to South Africa. Bulletin of the Atomic Scientists, 42(3):10, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ETI

[Ano86r] Anonymous. Earlier treaty interpretations. Bulletin of the Atomic Scientists, 42(1):33, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:EAC

[Ano86s] Anonymous. Erratum: [“Arms control: necessary process”]. Bulletin of the Atomic Scientists, 42(5):48, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pan86b].

Anonymous:1986:EGP

[Ano86t] Anonymous. Erratum: [Gorbachev’s policy innovations]. Bulletin of the Atomic Scientists, 42(1):57, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro85a].

Anonymous:1986:EWH

[Ano86u] Anonymous. Erratum: [‘What happened at Reactor Four’]. Bulletin of the Atomic Scientists, 42(10):52, December 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tho86a].

Anonymous:1986:ELG

[Ano86v] Anonymous. Excerpt from a letter from George Schultz to Barry Goldwater, read before the U.S. Senate on August 7, 1986. Bulletin of the Atomic Scientists, 42(9):30, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ENS

[Ano86w] Anonymous. Excerpt from NRDC–Soviet Academy of Sciences monitoring agreement. Bulletin of the Atomic Scientists, 42(8): 10, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ESA

[Ano86x] Anonymous. From the Editors: A security agenda. Bulletin of the Atomic Scientists, 42(6):2, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ECC

[Ano86y] Anonymous. From the Editors: Chernobyl in context. Bulletin of the Atomic Scientists, 42(7):2, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:EOM

[Ano86z] Anonymous. From the Editors: Orwell’s memory holes. Bulletin of the Atomic Scientists, 42(2):2, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:EPS

[Ano86-27] Anonymous. From the Editors: Population and security. Bulletin of the Atomic Scientists, 42(4):2, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ETV

[Ano86-28] Anonymous. From the Editors: Technological vulnerability. Bulletin of the Atomic Scientists, 42(3):2, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ETS

[Ano86-29] Anonymous. From the Editors: Testing and spending. Bulletin of the Atomic Scientists, 42(8):2, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:ETB

[Ano86-30] Anonymous. From the Editors: The test ban and Olof Palme. Bulletin of the Atomic Scientists, 42(5):2, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:EUS

[Ano86-31] Anonymous. From the Editors: U.S.–Soviet priorities. Bulletin of the Atomic Scientists, 42(9):2, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:EXR

[Ano86-32] Anonymous. From the Editors: X-ray laser and secrecy. Bulletin of the Atomic Scientists, 42(1):2, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:GRT

[Ano86-33] Anonymous. Glossary of radiation terms. Bulletin of the Atomic Scientists, 42(7):12, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:GWT

[Ano86-34] Anonymous. Glossary on weapon terminology. Bulletin of the Atomic Scientists, 42(10):9, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:LIF

[Ano86-35] Anonymous. Letter: Iceland — failure and hope. Bulletin of the Atomic Scientists, 42(10):2, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:PSW

[Ano86-36] Anonymous. Pugwash on space weapons. Bulletin of the Atomic Scientists, 42(4):46–47, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1986:WI

[Ano86-37] Anonymous. What is the IAEA? Bulletin of the Atomic Scientists, 42(6):29, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Aa

[Ano87a] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(1):60, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Ab

[Ano87b] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(2):64, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Ac

[Ano87c] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(2):64, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Ad

[Ano87d] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(3):53, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Ae

[Ano87e] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(6):62, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Af

[Ano87f] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(7):61, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Ag

[Ano87g] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(8):54, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:Ah

[Ano87h] Anonymous. Announcements. Bulletin of the Atomic Scientists, 43(10):50, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:ARS

[Ano87i] Anonymous. Arms race at sea: The Defense Department’s view of Soviet maritime strategy. Bulletin of the Atomic Scientists, 43(7):20, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:ACN

[Ano87j] Anonymous. Average cost of new U.S. nuclear power plant. Bulletin of the Atomic Scientists, 43(6):28, July/August 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BWW

[Ano87k] Anonymous. Biological weapons and what they do. Bulletin of the Atomic Scientists, 43(1):44, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBf

[Ano87l] Anonymous. Book note: Securing Europe’s Future, edited by Stephen J. Flanagan. Bulletin of the Atomic Scientists, 43(8): 52, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBe

[Ano87m] Anonymous. Book note: Strategic Defenses. Bulletin of the Atomic Scientists, 43(7):57, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBd

[Ano87n] Anonymous. Book notes: A Modern Day Yankee in a Connecticut Court, by Alan Lightman; U.S. Arms Exports: Policies and Contractors, by Paul L. Ferrari, Jeffrey W. Knopf, and Raúl L. Madrid; Science and Society: A Directory to Information Sources, Volume 1, Defence and Arms Control, by Canadian Student Pugwash. Bulletin of the Atomic Scientists, 43(6):58, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBc

[Ano87o] Anonymous. Book notes: Nuclear Terrorism: Defining the Threat, edited by Paul Leventhal and Yonah Alexander. Bulletin of the Atomic Scientists, 43(3):50, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBa

[Ano87p] Anonymous. Book notes: Psychology and the Prevention of Nuclear War, edited by Ralph K. White. Bulletin of the Atomic Scientists, 43(3):50, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBg

[Ano87q] Anonymous. Book notes: Strategic Antisubmarine Warfare and Naval Strategy, by Tom Stefanick; Soviet Nuclear Weapons Policy: A Research and Bibliographic Guide, by William C. Green. Bulletin of the Atomic Scientists, 43(9):50, November 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BNBb

[Ano87r] Anonymous. Book notes: World Weapon Database: Volume I: Soviet Missiles, by Baron Wright. Bulletin of the Atomic Scientists, 43(3):50, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BRB

[Ano87s] Anonymous. Book review: The Language of Nuclear War: An Intelligent Citizen’s Dictionary, by Eric Semler, James Benjamin, and Adam Cross. Bulletin of the Atomic Scientists, 43(5):52, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BAS

[Ano87t] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 43, 1987. Bulletin of the Atomic Scientists, 43(10):53–56, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:BWO

[Ano87u] Anonymous. Bulletin wins Olive Branch. Bulletin of the Atomic Scientists, 43(4):55, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:CAT

[Ano87v] Anonymous. Channels for arms transfers. Bulletin of the Atomic Scientists, 43(6):24, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:CDA

[Ano87w] Anonymous. Civil defense according to FEMA. Bulletin of the Atomic Scientists, 43(4):44, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:EFD

[Ano87x] Anonymous. Erratum: [four decades of irrationality]. Bulletin of the Atomic Scientists, 43(10):50, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shu87].

Anonymous:1987:FBW

[Ano87y] Anonymous. Fighting biological weapons research in the courts. Bulletin of the Atomic Scientists, 43(1):45, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:EBW

[Ano87z] Anonymous. From the Editors: Bulletin wins top award. Bulletin of the Atomic Scientists, 43(5):2, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:EMW

[Ano87-27] Anonymous. From the Editors: McCarran–Walter revisited. Bulletin of the Atomic Scientists, 43(4):2, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:HDP

[Ano87-28] Anonymous. The Hamburg disarmament proposals. Bulletin of the Atomic Scientists, 43(1):52, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:IWC

[Ano87-29] Anonymous. IPPNW World Congress. Bulletin of the Atomic Scientists, 43(4):55, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:MUA

[Ano87-30] Anonymous. Major U.S. arms sales offers, fiscal year 1986. Bulletin of the Atomic Scientists, 43(6):25, July/August 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:NVW

[Ano87-31] Anonymous. NATO vs. Warsaw Pact: A selected bean count. Bulletin of the Atomic Scientists, 43(10):13, December 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:NPB

[Ano87-32] Anonymous. Nobel Prize for Bulletin sponsor: [John C. Polanyi, Yuan T. Lee, and Dudley Herschbach, in Chemistry]. Bulletin of the Atomic Scientists, 43(1):59–60, January/February 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:PCS

[Ano87-33] Anonymous. Pugwash on conventional strategy. Bulletin of the Atomic Scientists, 43(10):14, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:REC

[Ano87-34] Anonymous. Reagan’s easy credit terms. Bulletin of the Atomic Scientists, 43(6):22, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:RSI

[Ano87-35] Anonymous. “Revealed: the secrets of Israel’s nuclear arsenal”. Bulletin of the Atomic Scientists, 43(2):17, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:RMFa

[Ano87-36] Anonymous. Ruth Mott Fund grant. Bulletin of the Atomic Scientists, 43(3):53, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:RMFb

[Ano87-37] Anonymous. Ruth Mott Fund grant. Bulletin of the Atomic Scientists, 43(4):55, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:SHE

[Ano87-38] Anonymous. Skirting human error: the Navy’s missile launch system. Bulletin of the Atomic Scientists, 43(1):38, January/ February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:SAR

[Ano87-39] Anonymous. Superpower arms race at sea. Bulletin of the Atomic Scientists, 43(7):13, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:TLL

[Ano87-40] Anonymous. Teller and the laser lobby. Bulletin of the Atomic Scientists, 43(8):21, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:TA

[Ano87-41] Anonymous. The Toshiba affair. Bulletin of the Atomic Scientists, 43(7):42, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:TUM

[Ano87-42] Anonymous. Trends in U.S. military assistance, 1981 to 1986. Bulletin of the Atomic Scientists, 43(6):23, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:USS

[Ano87-43] Anonymous. U.S. and Soviet strategic nuclear forces, 1980–1986. Bulletin of the Atomic Scientists, 43(4):57, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:UIV

[Ano87-44] Anonymous. The U.S. INF verification package. Bulletin of the Atomic Scientists, 43(5):15, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1987:UNW

[Ano87-45] Anonymous. U.S. nuclear weapons stockpile (June 1987). Bulletin of the Atomic Scientists, 43(5):56, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:Aa

[Ano88a] Anonymous. Announcements. Bulletin of the Atomic Scientists, 44(3):58, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:Ab

[Ano88b] Anonymous. Announcements. Bulletin of the Atomic Scientists, 44(9):46, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:BNBc

[Ano88c] Anonymous. Book note: Soviet Submarine Operations in Swedish Water: 1980–1986, by Milton Leitenberg. Bulletin of the Atomic Scientists, 44(9):44, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lei89b].

Anonymous:1988:BNBa

[Ano88d] Anonymous. Book note: World Weapon Database, Volume II: Soviet Military Aircraft, by Neta Crawford. Bulletin of the Atomic Scientists, 44(7):58, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:BNBb

[Ano88e] Anonymous. Book notes: Balancing the National Interest: U.S. National Security Export Controls and Global Economic Competition, by the National Academy of Sciences; Comprehensive Peace Education: Educating for Global Responsibility, by Betty A. Reardon; Educating for Global Responsibility: Teacher-Designed Curricula for Peace Education, K–12 edited by Betty A. Reardon; Nuclear Weapon Tests: Prohibition or Limitation?, edited by Jozef Goldblat and David Cox. Bulletin of the Atomic Scientists, 44(8):50–51, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:BAS

[Ano88f] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 44, 1988. Bulletin of the Atomic Scientists, 44(10):53–55, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:CVO

[Ano88g] Anonymous. Censored version of OTA report says SDI might work, a little bit. Bulletin of the Atomic Scientists, 44(9):15– 16, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:CWG

[Ano88h] Anonymous. A conventional warfare glossary. Bulletin of the Atomic Scientists, 44(7):26, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:CC

[Ano88i] Anonymous. The costs of cleaning up. Bulletin of the Atomic Scientists, 44(1):36, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:ES

[Ano88j] Anonymous. Ezra Sensibar, 1905–1988. Bulletin of the Atomic Scientists, 44(8):6, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:HGF

[Ano88k] Anonymous. The hierarchy of ground-force units. Bulletin of the Atomic Scientists, 44(7):27, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:HAW

[Ano88l] Anonymous. How AVLIS works. Bulletin of the Atomic Scientists, 44(9):29, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:R

[Ano88m] Anonymous. I. I. Rabi, 1898–1988. Bulletin of the Atomic Scientists, 44(2):8, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:JDC

[Ano88n] Anonymous. The Justice Department’s catch-22. Bulletin of the Atomic Scientists, 44(1):32, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:NDN

[Ano88o] Anonymous. Nonoffensive defense in the news. Bulletin of the Atomic Scientists, 44(7):15, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:OLS

[Ano88p] Anonymous. Options for limiting sea-launched cruise missiles. Bulletin of the Atomic Scientists, 44(9):40, November 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:PWC

[Ano88q] Anonymous. Pugwash on weapons cuts. Bulletin of the Atomic Scientists, 44(6):46, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:RMF

[Ano88r] Anonymous. Ruth Mott Fund grant. Bulletin of the Atomic Scientists, 44(1):54, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:SSS

[Ano88s] Anonymous. Soviet strength in science compared to U.S. strength. Bulletin of the Atomic Scientists, 44(8):30, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:TC

[Ano88t] Anonymous. The troublesome cruise. Bulletin of the Atomic Scientists, 44(9):38, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:UCP

[Ano88u] Anonymous. A united call for pacifist rights. Bulletin of the Atomic Scientists, 44(4):22, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:WID

[Ano88v] Anonymous. What IAEA did in one year. Bulletin of the Atomic Scientists, 44(10):16, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:WAW

[Ano88w] Anonymous. Who’s attacking whom? Bulletin of the Atomic Scientists, 44(7):36, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:WMS

[Ano88x] Anonymous. William M. Swartz, 1912–1987. Bulletin of the Atomic Scientists, 44(1):42, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1988:WS

[Ano88y] Anonymous. William Swartz, 1912–1987. Bulletin of the Atomic Scientists, 44(5):0, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:A

[Ano89a] Anonymous. Announcements. Bulletin of the Atomic Scientists, 45(4):55, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:ANJ

[Ano89b] Anonymous. Announcements: [new journal, Interdisciplinary Peace Research]. Bulletin of the Atomic Scientists, 45(8):45, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:ACC

[Ano89c] Anonymous. Arms control criteria: Arms control treaties and agreements since World War II. Bulletin of the Atomic Scientists, 45(4):36, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BNBd

[Ano89d] Anonymous. Book note: Arms Control and National Security: An Introduction. Bulletin of the Atomic Scientists, 45(10):48, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BNBc

[Ano89e] Anonymous. Book note: U.S. Arms Exports: Policies and Contractors, 1988 Edition, by Paul L. Ferrari, Raul L. Madrid, and Jeff Knopf. Bulletin of the Atomic Scientists, 45(6):48, July/ August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BNBa

[Ano89f] Anonymous. Book notes: Nuclear Heartland, by Samuel H. Day, Jr.; Women’s Foreign Policy Council Directory, by Mim Kelber. Bulletin of the Atomic Scientists, 45(1):63, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BNBb

[Ano89g] Anonymous. Book notes: The Zero Option: INF, West Germany, and Arms Control, by Thomas Risse-Kappen; The ACCESS Resource Guide: An International Directory of Information on War, Peace, and Security. Bulletin of the Atomic Scientists, 45(5):43, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BAS

[Ano89h] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 45, 1989. Bulletin of the Atomic Scientists, 45(10):53–56, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BBV

[Ano89i] Anonymous. Bulletin: Budget votes: neither yea nor nay. Bulletin of the Atomic Scientists, 45(8):4–5, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BDC

[Ano89j] Anonymous. Bulletin: Dove Creek, accident GWEN spots? Bulletin of the Atomic Scientists, 45(9):5, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BHR

[Ano89k] Anonymous. Bulletin: Hands off those rocks. Bulletin of the Atomic Scientists, 45(9):4–5, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BRW

[Ano89l] Anonymous. Bulletin: Renewable Wackersdorf. Bulletin of the Atomic Scientists, 45(8):7, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:BYB

[Ano89m] Anonymous. Bulletin: Youngest bomb father. Bulletin of the Atomic Scientists, 45(9):5, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:C

[Ano89n] Anonymous. Correction. Bulletin of the Atomic Scientists, 45(1): 63, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:DS

[Ano89o] Anonymous. Dangerous simulations. Bulletin of the Atomic Scientists, 45(3):17, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:FRD

[Ano89p] Anonymous. FAS recognizes DeWitt. Bulletin of the Atomic Scientists, 45(4):55, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:FYA

[Ano89q] Anonymous. Forty years ago in the Bulletin. Bulletin of the Atomic Scientists, 45(8):7, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:HPH

[Ano89r] Anonymous. How Point Hope, Alaska took on Edward Teller. Bulletin of the Atomic Scientists, 45(10):[cover], December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:JO

[Ano89s] Anonymous. Just an oversight. Bulletin of the Atomic Scientists, 45(6):12, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:LM

[Ano89t] Anonymous. Looking for markets. Bulletin of the Atomic Scientists, 45(5):29, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:LSW

[Ano89u] Anonymous. Lost at sea 1945–89: 50 weapons, 11 reactors. Bulletin of the Atomic Scientists, 45(6):22, July/August 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:MBD

[Ano89v] Anonymous. A message from the Board of Directors. Bulletin of the Atomic Scientists, 45(8):3, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:NIJ

[Ano89w] Anonymous. A new international journal: [Science and Global Security]. Bulletin of the Atomic Scientists, 45(4):12, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:RMF

[Ano89x] Anonymous. Ruth Mott Fund grant. Bulletin of the Atomic Scientists, 45(1):66, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:TYA

[Ano89y] Anonymous. Thirty years ago in the Bulletin. Bulletin of the Atomic Scientists, 45(9):7, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:TFY

[Ano89z] Anonymous. Twenty-five years ago in the Bulletin. Bulletin of the Atomic Scientists, 45(10):5, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:WNC

[Ano89-27] Anonymous. What’s next, coveralls? Bulletin of the Atomic Scientists, 45(8):7, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:WND

[Ano89-28] Anonymous. Why not now? Debating a nuclear-free millennium. Bulletin of the Atomic Scientists, 45(6):25, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Che89].

Anonymous:1990:Cb

[Ano90a] Anonymous. 140 and counting. Bulletin of the Atomic Scientists, 46(6):26, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABf

[Ano90b] Anonymous. 15 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(9):5, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABd

[Ano90c] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(7):5, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABg

[Ano90d] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(10):5, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABa

[Ano90e] Anonymous. 35 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(4):5, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABb

[Ano90f] Anonymous. 35 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(5):5, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABe

[Ano90g] Anonymous. 35 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(8):5, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YABc

[Ano90h] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(6):5, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:YBC

[Ano90i] Anonymous. 43 years of the Bulletin clock: a history of the Cold War. Bulletin of the Atomic Scientists, 46(3):58, April 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:APH

[Ano90j] Anonymous. Announcement: [photograph of Hiroshima]. Bulletin of the Atomic Scientists, 46(7):47, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:Aa

[Ano90k] Anonymous. Announcements. Bulletin of the Atomic Scientists, 46(1):47, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:Ab

[Ano90l] Anonymous. Announcements. Bulletin of the Atomic Scientists, 46(9):45, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNBc

[Ano90m] Anonymous. Book note: The World in Conflict 1989: War Annual 3, by John Laffin. Bulletin of the Atomic Scientists, 46(8): 44, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNBb

[Ano90n] Anonymous. Book note: Toward a Comprehensive Test Ban, by Steve Fetter. Bulletin of the Atomic Scientists, 46(4):50, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNBd

[Ano90o] Anonymous. Book notes: The Soviet Empire: Its Nations Speak Out, by Oleg Glebov and John Crowfoot; Voices of Glasnost, edited by Stephen F. Cohen and Katrina vanden Heuvel; The Global Ecology Handbook: What You Can Do About the Environmental Crisis. Bulletin of the Atomic Scientists, 46(9):44–45, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNBa

[Ano90p] Anonymous. Book notes: World Social and Military Expenditures 1989, by Ruth Leger Sivard; Superpowers at Sea: An Assessment of the Naval Arms Race, by Richard Fieldhouse and Shunji Taoka. Bulletin of the Atomic Scientists, 46(2):43, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BAS

[Ano90q] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 46, 1990. Bulletin of the Atomic Scientists, 46(10):52–56, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BAC

[Ano90r] Anonymous. Bulletin: Arms controllers v. Greenpeace. Bulletin of the Atomic Scientists, 46(9):4–6, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBL

[Ano90s] Anonymous. Bulletin: Biotech lab recalls biowar. Bulletin of the Atomic Scientists, 46(1):6, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBMa

[Ano90t] Anonymous. Bulletin: Bring in the marines. Bulletin of the Atomic Scientists, 46(6):4, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pow90b].

Anonymous:1990:BBS

[Ano90u] Anonymous. Bulletin: Bugs in space. Bulletin of the Atomic Scientists, 46(4):5, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BCT

[Ano90v] Anonymous. Bulletin: Cage those tigers. Bulletin of the Atomic Scientists, 46(9):6, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BC

[Ano90w] Anonymous. Bulletin: Catch-22. Bulletin of the Atomic Scientists, 46(7):6, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BCG

[Ano90x] Anonymous. Bulletin: Clean and green Stealth. Bulletin of the Atomic Scientists, 46(1):3, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BCP

[Ano90y] Anonymous. Bulletin: Coalition politics? Bulletin of the Atomic Scientists, 46(6):7, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BFS

[Ano90z] Anonymous. Bulletin: Further study needed. Bulletin of the Atomic Scientists, 46(6):4, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BGN

[Ano90-27] Anonymous. Bulletin: Gain a nation, lose a dump. Bulletin of the Atomic Scientists, 46(3):6, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BGT

[Ano90-28] Anonymous. Bulletin: Go tigers? Bulletin of the Atomic Scientists, 46(2):7–8, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BB

[Ano90-29] Anonymous. Bulletin: In brief. Bulletin of the Atomic Scientists, 46(2):6, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBWa

[Ano90-30] Anonymous. Bulletin: In brief: And the winner is; voice of Amerika?; cold hard cash; by any other name; home videoski. Bulletin of the Atomic Scientists, 46(4):4, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBD

[Ano90-31] Anonymous. Bulletin: In brief: Did he take a pill?; too cheap to meter; starry eyed [Edward Teller]; think no small thoughts; good timing?; no preconceptions. Bulletin of the Atomic Scientists, 46(7):4, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBJ

[Ano90-32] Anonymous. Bulletin: In brief: Just what we were afraid of; everything must go; what’s next?; new pals; runaround. Bulletin of the Atomic Scientists, 46(9):6, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBN

[Ano90-33] Anonymous. Bulletin: In brief: Now we’re getting somewhere; 10 millionth spelling error; where are they now?; what’s the hurry? Bulletin of the Atomic Scientists, 46(6):7, July/August 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBMb

[Ano90-34] Anonymous. Bulletin: In brief: The more things change …; whew! past the hard part; still a solution to global warming? Pebbles not so brilliant. Bulletin of the Atomic Scientists, 46 (8):6, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBWb

[Ano90-35] Anonymous. Bulletin: In brief: What about the doohickey?; Devil’s Island, 1990s-style; taming the T-55. Bulletin of the Atomic Scientists, 46(10):7, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BJC

[Ano90-36] Anonymous. Bulletin: Just call. Bulletin of the Atomic Scientists, 46(5):4, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ano90-45].

Anonymous:1990:BLO

[Ano90-37] Anonymous. Bulletin: Little orphan antitank missile. Bulletin of the Atomic Scientists, 46(8):6, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BMW

[Ano90-38] Anonymous. Bulletin: Mothball the weapons. Bulletin of the Atomic Scientists, 46(2):8, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNN

[Ano90-39] Anonymous. Bulletin: Next, nuclear-free weapons lab. Bulletin of the Atomic Scientists, 46(1):3–4, January/February 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNC

[Ano90-40] Anonymous. Bulletin: Noble case to court. Bulletin of the Atomic Scientists, 46(6):6–7, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BNW

[Ano90-41] Anonymous. Bulletin: Nuclear waste dumpsites. Bulletin of the Atomic Scientists, 46(7):6, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BOC

[Ano90-42] Anonymous. Bulletin: One crackdown leads to another. Bulletin of the Atomic Scientists, 46(1):3, January/February 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BPR

[Ano90-43] Anonymous. Bulletin: Press releases don’t tell all. Bulletin of the Atomic Scientists, 46(1):4–5, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BRS

[Ano90-44] Anonymous. Bulletin: Reserving space for war. Bulletin of the Atomic Scientists, 46(3):5, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BSS

[Ano90-45] Anonymous. Bulletin: Safe snooping. Bulletin of the Atomic Scientists, 46(8):5–6, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano90-36].

Anonymous:1990:BSW

[Ano90-46] Anonymous. Bulletin: Soviet wheeler-dealers. Bulletin of the Atomic Scientists, 46(3):6, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BGB

[Ano90-47] Anonymous. Bulletin: The good and bad news at Czechoslovak Life. Bulletin of the Atomic Scientists, 46(4):3, May 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BTT

[Ano90-48] Anonymous. Bulletin: Try, try again. Bulletin of the Atomic Scientists, 46(6):6, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BTB

[Ano90-49] Anonymous. Bulletin: Two bombs’ worth. Bulletin of the Atomic Scientists, 46(6):4, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BWD

[Ano90-50] Anonymous. Bulletin: Watch out DARPA. Bulletin of the Atomic Scientists, 46(9):4, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wol91b].

Anonymous:1990:BWM

[Ano90-51] Anonymous. Bulletin: Whose missiles are unsafe? Bulletin of the Atomic Scientists, 46(6):6, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:CCA

[Ano90-52] Anonymous. Correction: [count of Americans who have worked in nuclear arms production]. Bulletin of the Atomic Scientists, 46 (8):44, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:Ca

[Ano90-53] Anonymous. Corrections. Bulletin of the Atomic Scientists, 46 (2):47, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa89j, Lan89d, Lev90].

Anonymous:1990:CSA

[Ano90-54] Anonymous. Corrections: Stopping accidents after they’ve happened; what the fight is all about. Bulletin of the Atomic Scientists, 46(10):51, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fra90a, McC90].

Anonymous:1990:CSN

[Ano90-55] Anonymous. Countries that have not signed the NPT. Bulletin of the Atomic Scientists, 46(6):24–25, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:FYAa

[Ano90-56] Anonymous. Forty years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(1):5, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:FYAb

[Ano90-57] Anonymous. Forty years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(2):7, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:FNF

[Ano90-58] Anonymous. France and the nuclear free-for-all. Bulletin of the Atomic Scientists, 46(6):18, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:NG

[Ano90-59] Anonymous. NPT at a glance. Bulletin of the Atomic Scientists, 46(6):16, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Van90a].

Anonymous:1990:ND

[Ano90-60] Anonymous. The NPT on disarmament. Bulletin of the Atomic Scientists, 46(10):46, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:PKS

[Ano90-61] Anonymous. Peter Kapitsa: The scientist who talked back to Stalin. Bulletin of the Atomic Scientists, 46(3):26–33, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [dW90a].

Anonymous:1990:TFY

[Ano90-62] Anonymous. Twenty-five years ago in the Bulletin. Bulletin of the Atomic Scientists, 46(3):6, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:WCD

[Ano90-63] Anonymous. Watkins and critics disagree. Bulletin of the Atomic Scientists, 46(10):4, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:WHW

[Ano90-64] Anonymous. Weapons: How and what the Third World produced, 1982–1987. Bulletin of the Atomic Scientists, 46(4):21, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABh

[Ano91a] Anonymous. 10 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(8):5, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABd

[Ano91b] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(4):5, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABg

[Ano91c] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(7):5, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABc

[Ano91d] Anonymous. 35 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(3):5, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABa

[Ano91e] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(1):5, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABb

[Ano91f] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(2):5, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABe

[Ano91g] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(5):5, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABf

[Ano91h] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(6):5, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABi

[Ano91i] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(9):5, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YBCa

[Ano91j] Anonymous. 44 years of the Bulletin clock: a history of the Cold War. Bulletin of the Atomic Scientists, 47(10):50, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YBCb

[Ano91k] Anonymous. 44 years of the Bulletin Clock: A history of the Cold War. Bulletin of the Atomic Scientists, 47(10):50, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:YABj

[Ano91l] Anonymous. 45 years ago in the Bulletin. Bulletin of the Atomic Scientists, 47(10):5, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:AW

[Ano91m] Anonymous. Acronym watch. Bulletin of the Atomic Scientists, 47(8):6, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:AC

[Ano91n] Anonymous. An agenda for Congress. Bulletin of the Atomic Scientists, 47(8):37, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:A

[Ano91o] Anonymous. Announcements. Bulletin of the Atomic Scientists, 47(5):42, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BSU

[Ano91p] Anonymous. Bait and switch universal? Bulletin of the Atomic Scientists, 47(9):6, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BNBa

[Ano91q] Anonymous. Book note: Environmental Hazards of War: Releasing Dangerous Forces in an Industrialized World, by Arthur H. Westing. Bulletin of the Atomic Scientists, 47(3):46, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BNBc

[Ano91r] Anonymous. Book note: Physics and Nuclear Arms Today, edited by David Hafemeister. Bulletin of the Atomic Scientists, 47(6): 42–43, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BNBb

[Ano91s] Anonymous. Book notes: Preventing a Biological Arms Race, edited by Susan Wright; Nuclear Ambitions: The Spread of Nuclear Weapons 1989–1990, by Leonard S. Spector with Jacqueline R. Smith. Bulletin of the Atomic Scientists, 47(5):41–42, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BAS

[Ano91t] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 47, 1991. Bulletin of the Atomic Scientists, 47(10):45–48, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BWA

[Ano91u] Anonymous. Bulletin wins award [Joaquín Costa Award for Journalism from the Fundación Hogar del Empleado, Madrid, Spain]. Bulletin of the Atomic Scientists, 47(3):45, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BAH

[Ano91v] Anonymous. Bulletin: Attention hackers: Uncle Sam wants you. Bulletin of the Atomic Scientists, 47(10):5, December 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBT

[Ano91w] Anonymous. Bulletin: Bad timing. Bulletin of the Atomic Scientists, 47(4):4, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBBb

[Ano91x] Anonymous. Bulletin: Baubles from the bomb lab. Bulletin of the Atomic Scientists, 47(8):4–5, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBS

[Ano91y] Anonymous. Bulletin: Britain still ducking cleanup bill. Bulletin of the Atomic Scientists, 47(9):7, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BDCb

[Ano91z] Anonymous. Bulletin: Damage control. Bulletin of the Atomic Scientists, 47(5):4–5, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BDCa

[Ano91-27] Anonymous. Bulletin: Don’t have a cow, man. Bulletin of the Atomic Scientists, 47(3):5, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BEN

[Ano91-28] Anonymous. Bulletin: Energy’s “new culture” in action. Bulletin of the Atomic Scientists, 47(5):7, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BFS

[Ano91-29] Anonymous. Bulletin: FBI scrambles to stop scrambling. Bulletin of the Atomic Scientists, 47(6):4, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BGL

[Ano91-30] Anonymous. Bulletin: German largesse. Bulletin of the Atomic Scientists, 47(6):4–5, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BTD

[Ano91-31] Anonymous. Bulletin: Had taps, didn’t listen. Bulletin of the Atomic Scientists, 47(8):4, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BHE

[Ano91-32] Anonymous. Bulletin: Health and Energy agree, PSR disagrees. Bulletin of the Atomic Scientists, 47(3):6, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBA

[Ano91-33] Anonymous. Bulletin: In brief: Aloha; so what’s a neutron, anyway? Vice Admiral sinks the press; like being dropped from Who’s Who?; acronym watch. Bulletin of the Atomic Scientists, 47(8):6, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBBa

[Ano91-34] Anonymous. Bulletin: In brief: Best NATO buddies; slow but steady; censorship and Chernobyl; FOIA gets no report; more points of light. Bulletin of the Atomic Scientists, 47(6):6, July/ August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBC

[Ano91-35] Anonymous. Bulletin: In brief: Coke or Pepsi?; from black tulips to black diamonds; war games, one; war games, two. Bulletin of the Atomic Scientists, 47(10):6, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBD

[Ano91-36] Anonymous. Bulletin: In brief: Did my agent call?; running cold on warming; new fitness rationale?; just don’t use the nword; Lease-a-MiG? Bulletin of the Atomic Scientists, 47(4):6, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBG

[Ano91-37] Anonymous. Bulletin: In brief: Green-eyeshade jerks; how many secrets? that’s classified; silver lining — or maybe gold; waste not, want not? Bulletin of the Atomic Scientists, 47(5):6, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBI

[Ano91-38] Anonymous. Bulletin: In brief: Iraq helps Watkins; where are they now? Bulletin of the Atomic Scientists, 47(1):6, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBM

[Ano91-39] Anonymous. Bulletin: In brief: McCastro’s; risky communication; which report did they read?; seeking refuge in Berlin. Bulletin of the Atomic Scientists, 47(3):4, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBNa

[Ano91-40] Anonymous. Bulletin: In brief: Name game; “Do your part, drive smart” military-style; McMoney; pity the People’s Deputies; cheaper to move backyard. Bulletin of the Atomic Scientists, 47(2):4, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBNb

[Ano91-41] Anonymous. Bulletin: In brief: New, more open Energy; Navy’s video blues; bait and switch universal?; just saving stamps? Bulletin of the Atomic Scientists, 47(9):6, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BBR

[Ano91-42] Anonymous. Bulletin: In brief: Radiation game; Canadian defense dollars at work; U.S. defense dollars at work; recycled patriotism; who does his PR? Bulletin of the Atomic Scientists, 47(7): 4, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BSE

[Ano91-43] Anonymous. Bulletin: In search of enemies. Bulletin of the Atomic Scientists, 47(9):7, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BMDb

[Ano91-44] Anonymous. Bulletin: Male democracies. Bulletin of the Atomic Scientists, 47(1):6, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BMR

[Ano91-45] Anonymous. Bulletin: Medical research classified. Bulletin of the Atomic Scientists, 47(1):7, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BMDa

[Ano91-46] Anonymous. Bulletin: Miners, downwinders compensated. Bulletin of the Atomic Scientists, 47(1):5, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BNS

[Ano91-47] Anonymous. Bulletin: Nippon Senso buys American. Bulletin of the Atomic Scientists, 47(1):7, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BW

[Ano91-48] Anonymous. Bulletin: Not to worry. Bulletin of the Atomic Scientists, 47(7):4–5, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BPR

[Ano91-49] Anonymous. Bulletin: Pentagon’s rigid position on positioning. Bulletin of the Atomic Scientists, 47(7):3–4, September 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BPH

[Ano91-50] Anonymous. Bulletin: Preparedness at home. Bulletin of the Atomic Scientists, 47(3):5, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BPP

[Ano91-51] Anonymous. Bulletin: Prospering from the Polish press. Bulletin of the Atomic Scientists, 47(7):5, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BPL

[Ano91-52] Anonymous. Bulletin: PSR launches magazine. Bulletin of the Atomic Scientists, 47(6):5–6, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BRR

[Ano91-53] Anonymous. Bulletin: Rad records red hot. Bulletin of the Atomic Scientists, 47(4):6, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BRD

[Ano91-54] Anonymous. Bulletin: Rainbow dodges draft. Bulletin of the Atomic Scientists, 47(1):4, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BRP

[Ano91-55] Anonymous. Bulletin: Recycling passé? Bulletin of the Atomic Scientists, 47(6):5, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BSN

[Ano91-56] Anonymous. Bulletin: Saying nyet to nuclear. Bulletin of the Atomic Scientists, 47(5):5, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BTP

[Ano91-57] Anonymous. Bulletin: Tigers, part two. Bulletin of the Atomic Scientists, 47(6):5, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BUD

[Ano91-58] Anonymous. Bulletin: Uranium dumping. Bulletin of the Atomic Scientists, 47(8):5–6, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BUW

[Ano91-59] Anonymous. Bulletin: U.S. wraps Soviet reactor in red tape. Bulletin of the Atomic Scientists, 47(5):7, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:BH

[Ano91-60] Anonymous. The burden of history. Bulletin of the Atomic Scientists, 47(6):33, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:CF

[Ano91-61] Anonymous. The contamination factory. Bulletin of the Atomic Scientists, 47(8):34–39, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:ENE

[Ano91-62] Anonymous. Editorial: A new era. Bulletin of the Atomic Scientists, 47(10):3, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:HBC

[Ano91-63] Anonymous. How Complex Cleanup was written. Bulletin of the Atomic Scientists, 47(8):35, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:IHW

[Ano91-64] Anonymous. Iraq helps Watkins. Bulletin of the Atomic Scientists, 47(1):6, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:NEC

[Ano91-65] Anonymous. A new era. Bulletin of the Atomic Scientists, 47(10): 3, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:NAS

[Ano91-66] Anonymous. Nuclear ambitions: The spread of nuclear weapons 1989-1990 (book). Bulletin of the Atomic Scientists, 47(5):42, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:RNW

[Ano91-67] Anonymous. Reflections on the new world order. Bulletin of the Atomic Scientists, 47(5):19, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:SSB

[Ano91-68] Anonymous. START: A START briefing book. Bulletin of the Atomic Scientists, 47(9):24–25, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:WCC

[Ano91-69] Anonymous. What Congress can do. Bulletin of the Atomic Scientists, 47(7):38, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABf

[Ano92a] Anonymous. 10 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(6):5, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABg

[Ano92b] Anonymous. 15 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(7):5, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABd

[Ano92c] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(4):7, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABe

[Ano92d] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(5):6, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABh

[Ano92e] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(8):5, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABi

[Ano92f] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(9):6, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABb

[Ano92g] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(2):5, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABj

[Ano92h] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(10):4, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABa

[Ano92i] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(1):5, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:YABc

[Ano92j] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 48(3):5, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BAS

[Ano92k] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 48, 1992. Bulletin of the Atomic Scientists, 48(10):53–56, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BWO

[Ano92l] Anonymous. Bulletin wins Olive Branch. Bulletin of the Atomic Scientists, 48(7):7, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BAF

[Ano92m] Anonymous. Bulletin: Air Force fine tunes fighter info. Bulletin of the Atomic Scientists, 48(1):4, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BA

[Ano92n] Anonymous. Bulletin: Announcement. Bulletin of the Atomic Scientists, 48(8):5, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BAP

[Ano92o] Anonymous. Bulletin: Aspin’s 2 percent solution. Bulletin of the Atomic Scientists, 48(4):4–5, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBSa

[Ano92p] Anonymous. Bulletin: Business school for Red Army. Bulletin of the Atomic Scientists, 48(1):5, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BTN

[Ano92q] Anonymous. Bulletin: Two nuclear priesthoods?; and a taxing 12 minutes it is …; word of the year; nutty as a …; oh it’s too hot out, anyway; glow little Glomar. Bulletin of the Atomic Scientists, 48(2):6, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBC

[Ano92r] Anonymous. Bulletin: In brief: Closing the gap; if Bush — or Clinton — were really like Truman; safety first; the children’s story; if you can’t give it away, sell it; enriched through uranium. Bulletin of the Atomic Scientists, 48(9):4, November 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBF

[Ano92s] Anonymous. Bulletin: In brief: Forget earthquakes, let’s talk volcanology; a new dorm, maybe?; be prepared; ever vigilant. Bulletin of the Atomic Scientists, 48(8):4, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBH

[Ano92t] Anonymous. Bulletin: In brief: House notices end of Cold War …; NATO notices too; … pay any price; new recruits; un-word of the year; guns to soap. Bulletin of the Atomic Scientists, 48(7): 6, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBN

[Ano92u] Anonymous. Bulletin: In brief: New dissident; Gorby goes Hollywood?; oh capitalism!; location, location, location; comrade padre. Bulletin of the Atomic Scientists, 48(5):4, June 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBR

[Ano92v] Anonymous. Bulletin: In brief: Red faces; ESPying it; and what does Madonna think?; in the dark on “deep black”. Bulletin of the Atomic Scientists, 48(1):6, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBSb

[Ano92w] Anonymous. Bulletin: In brief: Spy to spy; the proud, the brave, the ready-to-leave; green bombs?; and when the sand runs out … ? Bulletin of the Atomic Scientists, 48(3):4, April 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBSc

[Ano92x] Anonymous. Bulletin: In brief: Stung by Stinger; more Russians = more spies?; glasnost à la Gates; vremia niu yorka v moskvie; is it a hobby, then? Bulletin of the Atomic Scientists, 48(6):6, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BBT

[Ano92y] Anonymous. Bulletin: In brief: Two stories; good try; futurology; roll call of shame; bidding war. Bulletin of the Atomic Scientists, 48(10):4, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BMO

[Ano92z] Anonymous. Bulletin: Missed opportunities. Bulletin of the Atomic Scientists, 48(1):5–6, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BMP

[Ano92-27] Anonymous. Bulletin: Missing pages [in September Bulletin]. Bulletin of the Atomic Scientists, 48(8):5, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BNE

[Ano92-28] Anonymous. Bulletin: New enemies in Pentagon basement. Bulletin of the Atomic Scientists, 48(2):7, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:BSN

[Ano92-29] Anonymous. Bulletin: Sinking nuclear ships. Bulletin of the Atomic Scientists, 48(2):4, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:CLO

[Ano92-30] Anonymous. Correction: [Less offense, more defense]. Bulletin of the Atomic Scientists, 48(4):46, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kre92a].

Anonymous:1992:CPU

[Ano92-31] Anonymous. Correction: [Proposed U.S. and C.I.S. strategic forces]. Bulletin of the Atomic Scientists, 48(5):49, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NA92i].

Anonymous:1992:EPP

[Ano92-32] Anonymous. Einstein Peace Prize awarded: [to Hans Bethe and Joseph Rotblat]. Bulletin of the Atomic Scientists, 48(10):5, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:KWC

[Ano92-33] Anonymous. Korean War chronology. Bulletin of the Atomic Scientists, 48(3):19, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:NMI

[Ano92-34] Anonymous. Notice: [misprinted issues with blank pages]. Bulletin of the Atomic Scientists, 48(4):47, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABa

[Ano93a] Anonymous. 10 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(1):4, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABf

[Ano93b] Anonymous. 10 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(6):7, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YALb

[Ano93c] Anonymous. 10 years ago …. Bulletin of the Atomic Scientists, 49(9):5, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABd

[Ano93d] Anonymous. 15 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(4):5, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABb

[Ano93e] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(2):5, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABc

[Ano93f] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(3):6, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YALa

[Ano93g] Anonymous. 30 years ago …. Bulletin of the Atomic Scientists, 49(8):4, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABe

[Ano93h] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(5):5, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YABg

[Ano93i] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 49(7):6, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:YALc

[Ano93j] Anonymous. 40 years ago …. Bulletin of the Atomic Scientists, 49(10):4, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:AC

[Ano93k] Anonymous. Aircraft carriers. Bulletin of the Atomic Scientists, 49(2):20, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:AFF

[Ano93l] Anonymous. Announcement: [“Facing the Fallout: Nuclear Weapons and the New World Disorder”]. Bulletin of the Atomic Scientists, 49(9):52, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:ARS

[Ano93m] Anonymous. Announcement: [Rebuilding security: The bomb, the debt, and the rainforest]. Bulletin of the Atomic Scientists, 49 (8):54, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BNB

[Ano93n] Anonymous. Book note: From Eros to Gaia, by Freeman Dyson. Bulletin of the Atomic Scientists, 49(5):54–55, June 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BAS

[Ano93o] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 49, 1993. Bulletin of the Atomic Scientists, 49(10):53–56, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBF

[Ano93p] Anonymous. Bulletin: Bernard T. Feld. Bulletin of the Atomic Scientists, 49(3):5, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BGD

[Ano93q] Anonymous. Bulletin: Glasnost à la the Defense Department. Bulletin of the Atomic Scientists, 49(7):6, September 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBC

[Ano93r] Anonymous. Bulletin: In brief: Clean bombs; we’re all connected; Ach du lieber; friends in high places; well, at least he’s consistent; just ask the neighbors. Bulletin of the Atomic Scientists, 49(6): 4, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBH

[Ano93s] Anonymous. Bulletin: In brief: How pyrrhic the victory?; crime pays — too much; searching the skies to proselytize; no time like the present; hazardous waste. Bulletin of the Atomic Scientists, 49(5):4, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBL

[Ano93t] Anonymous. Bulletin: In brief: Like manna from heaven; Woops!; but hey, its for peace; correction. Bulletin of the Atomic Scientists, 49(1):4, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBNb

[Ano93u] Anonymous. Bulletin: In brief: No, no secrets here; keeping the shirt on their backs; yes … yes, we do need to rationalize air; how many nuclear physicists does it take…; who’s the enemy? Bulletin of the Atomic Scientists, 49(10):5, December 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBNa

[Ano93v] Anonymous. Bulletin: In brief: ‘nuff said; book burning; peace and quiet; coming outen; tit for tat; Russian imperialism? Bulletin of the Atomic Scientists, 49(3):4, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBPb

[Ano93w] Anonymous. Bulletin: In brief: Prisoner of peace?; the way the wind is blowing; conversion capers; butter for guns?; blue dots indicate rest areas. Bulletin of the Atomic Scientists, 49(8):6, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBPa

[Ano93x] Anonymous. Bulletin: In brief: Pushing the button; my kingdom for a radio; ten-hut, war lovers; the 11 percent solution. Bulletin of the Atomic Scientists, 49(4):4, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBS

[Ano93y] Anonymous. Bulletin: In brief: Say again?; glasnost à la the Defense Department; brimming with redeeming social value; the last to know?; social realism, recycled. Bulletin of the Atomic Scientists, 49(7):6, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBO

[Ano93z] Anonymous. Bulletin: In brief: The old chiefs and Indians thing; did they send NASA a catalogue?; the winds of change; I’ll have the submarine sandwich. Bulletin of the Atomic Scientists, 49(9): 4, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BBW

[Ano93-27] Anonymous. Bulletin: In brief: The war that made the desert bloom; Tikho! (shush!); no artistic license; make that nine out of ten. Bulletin of the Atomic Scientists, 49(2):4, March 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sta93].

Anonymous:1993:BLW

[Ano93-28] Anonymous. Bulletin: Lederman wins Fermi Award. Bulletin of the Atomic Scientists, 49(7):7, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:BSR

[Ano93-29] Anonymous. Bulletin: Simpson recognized as space pioneer. Bulletin of the Atomic Scientists, 49(4):7, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:CTA

[Ano93-30] Anonymous. Correction: [Test anxiety]. Bulletin of the Atomic Scientists, 49(9):55, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kia93].

Anonymous:1993:CGB

[Ano93-31] Anonymous. Corrections: [German bomb program and Team B]. Bulletin of the Atomic Scientists, 49(4):6, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet93, CP93a].

Anonymous:1993:MWA

[Ano93-32] Anonymous. The men and women of the armed forces. Bulletin of the Atomic Scientists, 49(2):23, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:NA

[Ano93-33] Anonymous. Nuclear asymmetry. Bulletin of the Atomic Scientists, 49(2):14, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:RD

[Ano93-34] Anonymous. The real danger. Bulletin of the Atomic Scientists, 49(3):29, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:SPP

[Ano93-35] Anonymous. Secret [plutonium] production plants. Bulletin of the Atomic Scientists, 49(5):17, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:S

[Ano93-36] Anonymous. Submarines. Bulletin of the Atomic Scientists, 49 (2):21, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:TK

[Ano93-37] Anonymous. Tank killers. Bulletin of the Atomic Scientists, 49 (2):19, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1993:T

[Ano93-38] Anonymous. Tanks. Bulletin of the Atomic Scientists, 49(2):18, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YALa

[Ano94a] Anonymous. 10 years ago …. Bulletin of the Atomic Scientists, 50(1):8, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YALc

[Ano94b] Anonymous. 15 years ago …. Bulletin of the Atomic Scientists, 50(3):9, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YABb

[Ano94c] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 50(5):7, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YABc

[Ano94d] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 50(5):7, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YALb

[Ano94e] Anonymous. 30 years ago …. Bulletin of the Atomic Scientists, 50(2):8, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YALd

[Ano94f] Anonymous. 40 years ago …. Bulletin of the Atomic Scientists, 50(6):6, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:YABa

[Ano94g] Anonymous. 45 years ago in the Bulletin. Bulletin of the Atomic Scientists, 50(4):7, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BNB

[Ano94h] Anonymous. Book note: The Fermi Solution: Essays on Science, by Hans Christian von Baeyer. Bulletin of the Atomic Scientists, 50(1):61, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BAS

[Ano94i] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 50, 1994. Bulletin of the Atomic Scientists, 50(6):62–64, December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BBB

[Ano94j] Anonymous. Bulletin: In brief: Beer — it’s good for you; memorial reefs; life after death; netting the early bird; de-netting the Pentagon; Catch-22; doing a little bizniz; conversion capers; whistle while you … golf; promises, promises. Bulletin of the Atomic Scientists, 50(2):5–6, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BBG

[Ano94k] Anonymous. Bulletin: In brief: Go figure; don’t leave home without it; Communism as theme park; going bats; déjà vu all over again; in case you wondered; the Americanization of Natasha; fun and games; Aloha. Bulletin of the Atomic Scientists, 50(1): 5, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BBPb

[Ano94l] Anonymous. Bulletin: In brief: Is this the peace dividend? curiouser and curiouser; decommissioning — man and machine; a failure to communicate; welcome to the Ivan Hilton; immaculate incapsulation?; a tragic inability to delegate; no secrets. Bulletin of the Atomic Scientists, 50(6):7–8, November/December 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jag95].

Anonymous:1994:BBN

[Ano94m] Anonymous. Bulletin: In brief: Nobody told us; waste not, want not; military discipline; no open skies at CIA; feeling insecure lately?; more reason for panic; let the games begin; “Nuke the whales” agreement takes effect; conversion capers, continued; aged to perfection. Bulletin of the Atomic Scientists, 50(4):7–8, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BBPa

[Ano94n] Anonymous. Bulletin: In brief: Please pass the potatoes; maybe violence causes violence; Flipper denied honorable discharge; always the lucky one; hot air; cabinet of curiosities; oh … that’s all right, then; tit for tat; conversion capers, continued; just call MiGs Etcetera. Bulletin of the Atomic Scientists, 50(5):7– 8, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:BBR

[Ano94o] Anonymous. Bulletin: In brief: Remember Windscale?; 4,000 souvenirs; rapid-fire advice; bringing plastic to the people; giving it away …; … or selling it off; putting the terror in counterterrorism; really out of the loop. Bulletin of the Atomic Scientists, 50(3):5–6, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:CEE

[Ano94p] Anonymous. Correction: [Energy enters guilty plea]. Bulletin of the Atomic Scientists, 50(3):9, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mak94a].

Anonymous:1994:CPL

[Ano94q] Anonymous. Correction: [picture of Leo Szilard]. Bulletin of the Atomic Scientists, 50(5):3, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Les94].

Anonymous:1994:CWB

[Ano94r] Anonymous. Correction: [What a blast]. Bulletin of the Atomic Scientists, 50(1):7, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot93t].

Anonymous:1994:CNL

[Ano94s] Anonymous. Corrections: [Non-lethal weapons may violate treaties and Taking orders from Little Havana]. Bulletin of the Atomic Scientists, 50(6):60, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ros94, BT94].

Anonymous:1994:DCV

[Ano94t] Anonymous. Dollars for children v. dollars for guns. Bulletin of the Atomic Scientists, 50(2):7, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:PSC

[Ano94u] Anonymous. Plutonium stories: Cal-3, HP-9, HP-6, HP-3, Cal-1. Bulletin of the Atomic Scientists, 50(2):21–24, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:RBH

[Ano94v] Anonymous. Russia’s “blue helmets”. Bulletin of the Atomic Scientists, 50(1):58, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:TAA

[Ano94w] Anonymous. Tajikistan: Another Afghan war? Bulletin of the Atomic Scientists, 50(1):56–59, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:UEa

[Ano94x] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 50(2):8, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:UEb

[Ano94y] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 50(4):13, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:UEc

[Ano94z] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 50(5):9, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:UEd

[Ano94-27] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 50(6):7, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:UTV

[Ano94-28] Anonymous. Uzbekistan: Tamerlane v. Marx. Bulletin of the Atomic Scientists, 50(1):48–51, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:P

[Ano95a] Anonymous. The 20 principles. Bulletin of the Atomic Scientists, 51(4):29, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:YABb

[Ano95b] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 51(3):8, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:YABe

[Ano95c] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 51(6):9, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:YABd

[Ano95d] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 51(5):9, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:YA

[Ano95e] Anonymous. 40 years ago. Bulletin of the Atomic Scientists, 51(2):8, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:YABa

[Ano95f] Anonymous. 45 years ago in the Bulletin. Bulletin of the Atomic Scientists, 51(1):9, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:YABc

[Ano95g] Anonymous. 45 years ago in the Bulletin. Bulletin of the Atomic Scientists, 51(4):9, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BAS

[Ano95h] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 51, 1995. Bulletin of the Atomic Scientists, 51(6):77–80, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BBD

[Ano95i] Anonymous. Bulletin: In brief: Don’t phone home; Hitler’s hideaway as hot potato; Teller wants more; peace dividend update; secret — or dumb? counting (radiation-free) sheep; widening the customer base; bait and switch?; taking his word for it. Bulletin of the Atomic Scientists, 51(3):7–8, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BBG

[Ano95j] Anonymous. Bulletin: In brief: Getting one’s goat; blacking out the EPA; distinguished alumni; NIFty arrangement; don’t worry, be happy; fun with freebies; hope springs eternal; call your travel agent. Bulletin of the Atomic Scientists, 51(4):7–8, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BBL

[Ano95k] Anonymous. Bulletin: In brief: Look on the bright side; the new spirit of openness, sort of; special to the hacker hot line; circular thinking; less than meets the eye; embalming update. Bulletin of the Atomic Scientists, 51(2):7–8, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BBSa

[Ano95l] Anonymous. Bulletin: In brief: Sort of open; now that’s patented protection; nothing sacred; they never met a weapon they didn’t like; burrowing astronomers?; arms and no men; capers at Quad Cities. Bulletin of the Atomic Scientists, 51(1):7–8, January/ February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BBSb

[Ano95m] Anonymous. Bulletin: In brief: Surfin’ the what? no bodies on the razor wire?; Clinton’s plowshare; tit for tat, or not; between a rock and a hard place; cash-and-carry; but not one cent for openness; building public trust, part 27. Bulletin of the Atomic Scientists, 51(5):7–8, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:BBR

[Ano95n] Anonymous. Bulletin: In brief: The right to bear (nuclear) arms; with an 800,000-pound roof; wasn’t once enough? know thy neighbor; not enough gossip; a little help; let them eat frigates; nuclear power one, wave power zero; will test site let sun shine in?; and tasty, too. Bulletin of the Atomic Scientists, 51(6):7–8, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:CL

[Ano95o] Anonymous. Coming up …. Bulletin of the Atomic Scientists, 51(6):10, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:CSP

[Ano95p] Anonymous. Correction: [A sense of proportion]. Bulletin of the Atomic Scientists, 51(6):73, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano95v].

Anonymous:1995:ES

[Ano95q] Anonymous. Edward A. Shils. Bulletin of the Atomic Scientists, 51(3):6–7, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:EY

[Ano95r] Anonymous. Embarrassed? Yes. Bulletin of the Atomic Scientists, 51(4):6, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:EGE

[Ano95s] Anonymous. The Enola Gay Exhibit Advisory Board. Bulletin of the Atomic Scientists, 51(3):31, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:HDW

[Ano95t] Anonymous. How did we get in this mess? Bulletin of the Atomic Scientists, 51(3):36–41, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:IP

[Ano95u] Anonymous. Information, please. Bulletin of the Atomic Scientists, 51(5):51, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:SP

[Ano95v] Anonymous. A sense of proportion. Bulletin of the Atomic Scientists, 51(5):32–33, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano95p].

Anonymous:1995:SC

[Ano95w] Anonymous. Subramanyam Chandrasekhar. Bulletin of the Atomic Scientists, 51(6):10, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:TFP

[Ano95x] Anonymous. Time flies — to Peace Museum. Bulletin of the Atomic Scientists, 51(5):9, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:UEa

[Ano95y] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 51(1):7, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:UEb

[Ano95z] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 51(2):7, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:UEc

[Ano95-27] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 51(3):59, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1995:UEd

[Ano95-28] Anonymous. Upcoming events. Bulletin of the Atomic Scientists, 51(4):8, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:YABb

[Ano96a] Anonymous. 10 years ago in the Bulletin. Bulletin of the Atomic Scientists, 52(3):6, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:YABe

[Ano96b] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 52(6):9, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:YABa

[Ano96c] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 52(2):6, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:YABd

[Ano96d] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 52(5):7, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:YABc

[Ano96e] Anonymous. 45 years ago in the Bulletin. Bulletin of the Atomic Scientists, 52(4):7, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:AL

[Ano96f] Anonymous. Alexander Langsdorf. Bulletin of the Atomic Scientists, 52(4):10, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BS

[Ano96g] Anonymous. The board statement. Bulletin of the Atomic Scientists, 52(2):21, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BAS

[Ano96h] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 52, 1996. Bulletin of the Atomic Scientists, 52(6):68–72, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BIG

[Ano96i] Anonymous. Bulletin: Impasse in Geneva. Bulletin of the Atomic Scientists, 52(5):9, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BBW

[Ano96j] Anonymous. Bulletin: In brief: And the winner is …; knowing the enemy; reality mirrors art; green submarines; conversion capers, continued; a match made in heaven; guess who?; where the black stork and the capercailly play; foiling the snoops. Bulletin of the Atomic Scientists, 52(4):7–8, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BBI

[Ano96k] Anonymous. Bulletin: In brief: Ignoble; intolerable; et tu, Canada; conversion capers, cont.; clear-eyed in Warsaw; go, Trabi, go; camo chic. Bulletin of the Atomic Scientists, 52(1): 7–8, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [RL96].

Anonymous:1996:BBSa

[Ano96l] Anonymous. Bulletin: In brief: Set for life; better living through chemistry?; in the trenches; greenback disarmament; space invaders; “A great fixer-upper”; angles and loopholes. Bulletin of the Atomic Scientists, 52(2):7, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BBSb

[Ano96m] Anonymous. Bulletin: In brief: Spinning in his grave?; smiling through; O.K., so it’s not Jurassic Park; sheep may safely graze; don’t bank on it; from missiles to media; hope blooms; selective vision; never say die; for Los Alamos, it’s the pits. Bulletin of the Atomic Scientists, 52(3):7–8, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BBT

[Ano96n] Anonymous. Bulletin: In brief; titski for tatski; an ounce of prevention; spin cycle; nuclear underwear; jolly good oversight; “Russia — the pride is back”; no editors need apply; on the frontiers of the “Faustian bargain”. Bulletin of the Atomic Scientists, 52(6):7–8, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BBU

[Ano96o] Anonymous. Bulletin: In brief: Uh-oh; a better idea; of words and deeds; one of every three dollars…; and speaking of $800 billion…; values-free science; not the Swiss, too!; telling time tomorrow; making mea culpa pay. Bulletin of the Atomic Scientists, 52(5):7–8, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BNR

[Ano96p] Anonymous. Bulletin: Nouveaux Russians. Bulletin of the Atomic Scientists, 52(5):10, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:BWH

[Ano96q] Anonymous. Bulletin: Wanted: Historical insight. Bulletin of the Atomic Scientists, 52(1):8, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:CCS

[Ano96r] Anonymous. Chernobyl. Bulletin of the Atomic Scientists, 52 (3):20–21, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bru96].

Anonymous:1996:C

[Ano96s] Anonymous. The committee. Bulletin of the Atomic Scientists, 52(1):40, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:D

[Ano96t] Anonymous. The debate. Bulletin of the Atomic Scientists, 52 (2):44, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:E

[Ano96u] Anonymous. The experiments. Bulletin of the Atomic Scientists, 52(1):39, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:FSTa

[Ano96v] Anonymous. Faculty seminar on teaching strategic weapons proliferation. Bulletin of the Atomic Scientists, 52(3):19, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:FSTb

[Ano96w] Anonymous. Faculty seminar on teaching strategic weapons proliferation. Bulletin of the Atomic Scientists, 52(4):9, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:KR

[Ano96x] Anonymous. Key recommendations. Bulletin of the Atomic Scientists, 52(1):38, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:MI

[Ano96y] Anonymous. More information. Bulletin of the Atomic Scientists, 52(1):37, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:NLO

[Ano96z] Anonymous. New life for an old idea. Bulletin of the Atomic Scientists, 52(2):24–25, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:SCP

[Ano96-27] Anonymous. Stimson Center Project Steering Committee. Bulletin of the Atomic Scientists, 52(2):55, March/April 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:TYA

[Ano96-28] Anonymous. Ten years ago in the Bulletin. Bulletin of the Atomic Scientists, 52(1):7, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:Vb

[Ano96-29] Anonymous. Vitals. Bulletin of the Atomic Scientists, 52(6):41, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:Va

[Ano96-30] Anonymous. The vote. Bulletin of the Atomic Scientists, 52(5): 41, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:WCP

[Ano96-31] Anonymous. The World Court Project. Bulletin of the Atomic Scientists, 52(2):42, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:YABd

[Ano97a] Anonymous. 10 years ago in the Bulletin. Bulletin of the Atomic Scientists, 53(4):8, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:YABb

[Ano97b] Anonymous. 15 years ago in the Bulletin. Bulletin of the Atomic Scientists, 53(2):10, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:YABe

[Ano97c] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 53(5):11, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:YABf

[Ano97d] Anonymous. 20 years ago in the Bulletin. Bulletin of the Atomic Scientists, 53(6):6, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:YABa

[Ano97e] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 53(1):6, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:YABc

[Ano97f] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 53(3):6, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:AQ

[Ano97g] Anonymous. Author’s query. Bulletin of the Atomic Scientists, 53(6):63, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BAS

[Ano97h] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 53, 1997. Bulletin of the Atomic Scientists, 53(6):68–71, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BBB

[Ano97i] Anonymous. Bulletin: In brief: A brick of a car; … and an element to be named later; the birds and the bees; tourists trapped; sim war; are teraflops great or what?; no good deed goes unpunished; investing for war. Bulletin of the Atomic Scientists, 53(6):7–8, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BBO

[Ano97j] Anonymous. Bulletin: In brief: And for official state bird, the loon; a nuclear family, I presume?; singing for science; but is it art?; still batting a thousand; sex selection, the hard way; friends in high places. Bulletin of the Atomic Scientists, 53(4):7–8, July/ August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BBL

[Ano97k] Anonymous. Bulletin: In brief: But, like, it’s history, man; tectonic terror; a true zero-sum game; and if we had wings …; Star Wars forever; very generous of them; Toy Story, part two; get ’em while they’re hot. Bulletin of the Atomic Scientists, 53(1): 7–8, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BBM

[Ano97l] Anonymous. Bulletin: In brief: Moxing it up; demanding a recount; well said; better than advertised; defense conversion, part 27; defense conversion, part 28; desertification by design; blam! blam! Bulletin of the Atomic Scientists, 53(5):7–8, September/ October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Roc97b].

Anonymous:1997:BBP

[Ano97m] Anonymous. Bulletin: In brief: Post-Cold War capers; enemies list running low?; recycling; bang on the money; get real, part I; get real, part II; will they call it the UNRA?; comics for Congress; whither Energy. Bulletin of the Atomic Scientists, 53(2):7–8, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BBI

[Ano97n] Anonymous. Bulletin: In brief: The itty bitty bomb threat; for our own protection?; I spy; the granny v. grunt war; goodbye, Superphénix; back to nature; plutonium in a can; now that’s rent control. Bulletin of the Atomic Scientists, 53(3):7–8, May/ June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BT

[Ano97o] Anonymous. Bulletin: Thanks! Bulletin of the Atomic Scientists, 53(5):6, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:BBW

[Ano97p] Anonymous. Bulletin: The Bulletin welcomes George Lopez. Bulletin of the Atomic Scientists, 53(4):6, July/August 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:CGC

[Ano97q] Anonymous. Correction: [The ghost of Christmas past]. Bulletin of the Atomic Scientists, 53(6):62, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rof97].

Anonymous:1997:DCS

[Ano97r] Anonymous. The Deep Cuts Study Group. Bulletin of the Atomic Scientists, 53(3):37, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1997:SH

[Ano97s] Anonymous. Summit highlights. Bulletin of the Atomic Scientists, 53(4):29, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Another:1997:GR

[Ano97t] Another. Going on record. Bulletin of the Atomic Scientists, 53 (2):35, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YABd

[Ano98a] Anonymous. 15 years ago in the Bulletin. Bulletin of the Atomic Scientists, 54(4):8, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YABa

[Ano98b] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 54(1):8, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YABc

[Ano98c] Anonymous. 25 years ago in the Bulletin. Bulletin of the Atomic Scientists, 54(3):11, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YABf

[Ano98d] Anonymous. 30 years ago in the Bulletin. Bulletin of the Atomic Scientists, 54(6):11, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YABb

[Ano98e] Anonymous. 40 years ago in the Bulletin. Bulletin of the Atomic Scientists, 54(2):6, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YABe

[Ano98f] Anonymous. 50 years ago in the Bulletin. Bulletin of the Atomic Scientists, 54(5):8, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BNB

[Ano98g] Anonymous. Book note: Endless Frontier: Vannevar Bush, Engineer of the American Century, by G. Pascal Zachary. Bulletin of the Atomic Scientists, 54(4):64, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BRA

[Ano98h] Anonymous. Book reviews: Abolition primers: Nuclear Weapons: the Road to Zero, edited by Joseph Rotblat; The Gift of Time, by Jonathan Schell. Bulletin of the Atomic Scientists, 54(5):65, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BAS

[Ano98i] Anonymous. Bulletin of the Atomic Scientists: Index to Volume 54, 1998. Bulletin of the Atomic Scientists, 54(6):68–71, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BWK

[Ano98j] Anonymous. The Bulletin welcomes Kate Holder. Bulletin of the Atomic Scientists, 54(1):3, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BBW

[Ano98k] Anonymous. Bulletin: In brief: Are we there yet?; desktop publishing; lies, damned lies, and …; every day in every way …; domain dominion; dangerous zeroes; man v. nature; oops. Bulletin of the Atomic Scientists, 54(4):9–10, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BBD

[Ano98l] Anonymous. Bulletin: In brief: Don’t tell Cap Weinberger; lost forever, Clementine?; let ’em read comic books; next time, go for unlisted; but not below public opinion; meow; no falling sky yet; neighborly concern; science at play. Bulletin of the Atomic Scientists, 54(1):7–8, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BBL

[Ano98m] Anonymous. Bulletin: In brief: Like, totally rad; the other “atomic clock”; all quiet on the Western front; everyone’s going to want one; testing, testing, part 1; testing, testing, part 2; not the final frontier, after all. Bulletin of the Atomic Scientists, 54(5):7–8, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

www.boulder.nist.gov/timefreq/javaclck.html.

Anonymous:1998:BBM

[Ano98n] Anonymous. Bulletin: In brief: My dinner with Jehangir; if it’s not about Monica; anti-nuke camp; commie gastronomy; Uncle Big Brother?; if each author writes just three words … ? the talking cure. Bulletin of the Atomic Scientists, 54(6):7–8, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:BBS

[Ano98o] Anonymous. Bulletin: In brief: Semi-super computing; reviving the moldy oldies; going for the gold; cleaning up [on] the colonies; blame it on “concept stability”; one-time special offer, with coupon; believers not welcome; no one here by that name. Bulletin of the Atomic Scientists, 54(3):9–10, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:TDM

[Ano98p] Anonymous. Bulletin: In brief: Thanks, but don’t mention it to the Smithsonian; toys in the attic; don’t dis the broccoli; “Significant development challenges”; I see London, I see France; what, me worry?; who woulda thunk it? Bulletin of the Atomic Scientists, 54(2):7, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:LR

[Ano98q] Anonymous. Left or right. Bulletin of the Atomic Scientists, 54(3):29, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:SFT

[Ano98r] Anonymous. Strategic forces, today and tomorrow. Bulletin of the Atomic Scientists, 54(2):29, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:YS

[Ano98s] Anonymous. The youngest spy. Bulletin of the Atomic Scientists, 54(1):47, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:AES

[Ano99a] Anonymous. The agony and the ex-Stasi. Bulletin of the Atomic Scientists, 55(4):66–68, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/4/66.

Anonymous:1999:AV

[Ano99b] Anonymous. America the vulnerable? Bulletin of the Atomic Scientists, 55(1):72–74, January/February 1999. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/72.

Anonymous:1999:BT

[Ano99c] Anonymous. The B team. Bulletin of the Atomic Scientists, 55(1):70–71, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/1/70.

Anonymous:1999:BP

[Ano99d] Anonymous. Beyond the pale. Bulletin of the Atomic Scientists, 55(1):74–76, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/74.

Anonymous:1999:BNb

[Ano99e] Anonymous. A bio nightmare. Bulletin of the Atomic Scientists, 55(4):69–71, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/4/69.

Anonymous:1999:BNa

[Ano99f] Anonymous. Book note. Bulletin of the Atomic Scientists, 55(1): 76, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/1/76.

Anonymous:1999:Ba

[Ano99g] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 55(1): 5–9, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/1/5.

Anonymous:1999:Bc

[Ano99h] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 55(2):7– 11, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
2/7.

Anonymous:1999:Bd

[Ano99i] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 55 (3):7–12, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/3/7.

Anonymous:1999:Be

[Ano99j] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 55(4): 7–13, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/4/7.

Anonymous:1999:Bg

[Ano99k] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 55(5):7– 13, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/5/7.

Anonymous:1999:Bh

[Ano99l] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 55(6): 5–11, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/6/5.

Anonymous:1999:CNF

[Ano99m] Anonymous. Chinese nuclear forces, 1999. Bulletin of the Atomic Scientists, 55(3):79, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:FE

[Ano99n] Anonymous. A feckless embargo. Bulletin of the Atomic Scientists, 55(5):68–69, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/5/68.

Anonymous:1999:FBN

[Ano99o] Anonymous. French and British nuclear forces, 1999. Bulletin of the Atomic Scientists, 55(4):77, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:Bb

[Ano99p] Anonymous. In brief. Bulletin of the Atomic Scientists, 55(1): 7, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:Bf

[Ano99q] Anonymous. In brief. Bulletin of the Atomic Scientists, 55(4):9, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:Bi

[Ano99r] Anonymous. In brief. Bulletin of the Atomic Scientists, 55(6):7, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:Ib

[Ano99s] Anonymous. Index 1999. Bulletin of the Atomic Scientists, 55(6): 68–71, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/6/68.

Anonymous:1999:Ia

[Ano99t] Anonymous. Injected! Bulletin of the Atomic Scientists, 55(6): 58–61, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/6/58.

Anonymous:1999:La

[Ano99u] Anonymous. Letters. Bulletin of the Atomic Scientists, 55(1): ??, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/1/3.

Anonymous:1999:Lb

[Ano99v] Anonymous. Letters. Bulletin of the Atomic Scientists, 55(3):??, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
3/3.

Anonymous:1999:Lc

[Ano99w] Anonymous. Letters. Bulletin of the Atomic Scientists, 55 (4):??, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/4/3.

Anonymous:1999:Ld

[Ano99x] Anonymous. Letters. Bulletin of the Atomic Scientists, 55(5):3– 5, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/5/3.

Anonymous:1999:Le

[Ano99y] Anonymous. Letters. Bulletin of the Atomic Scientists, 55(6):??, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/6/3.

Anonymous:1999:PT

[Ano99z] Anonymous. The poll truth. Bulletin of the Atomic Scientists, 55 (6):62–63, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/6/62.

Anonymous:1999:RSN

[Ano99-27] Anonymous. Russian strategic nuclear forces, end of 1998. Bulletin of the Atomic Scientists, 55(2):62, March/April 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:U

[Ano99-28] Anonymous. Überterrorists. Bulletin of the Atomic Scientists, 55 (5):67–68, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/5/67.

Anonymous:1999:USN

[Ano99-29] Anonymous. U.S. strategic nuclear forces, end of 1998. Bulletin of the Atomic Scientists, 55(1):78, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1999:VWD

[Ano99-30] Anonymous. The volunteers who didn’t. Bulletin of the Atomic Scientists, 55(6):61–62, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/6/61.

Anonymous:1999:WH

[Ano99-31] Anonymous. Witch hunt. Bulletin of the Atomic Scientists, 55 (4):68–69, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/4/68.

Anonymous:2000:Ba

[Ano00a] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 56(1):7– 10, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/1/7.

Anonymous:2000:Bc

[Ano00b] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 56(2):6– 10, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
2/6.

Anonymous:2000:Bd

[Ano00c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 56 (3):6–11, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/3/6.

Anonymous:2000:Be

[Ano00d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 56(4): 6–12, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/4/6.

Anonymous:2000:Bf

[Ano00e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 56(5):4– 9, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/5/4.

Anonymous:2000:Bg

[Ano00f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 56(6): 7–12, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/6/7.

Anonymous:2000:Bb

[Ano00g] Anonymous. In brief. Bulletin of the Atomic Scientists, 56(1): 9, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2000:I

[Ano00h] Anonymous. Index 2000. Bulletin of the Atomic Scientists, 56(6): 80–83, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/6/80.

Anonymous:2000:La

[Ano00i] Anonymous. Letters. Bulletin of the Atomic Scientists, 56(1): 3–5, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/1/3.

Anonymous:2000:Lb

[Ano00j] Anonymous. Letters. Bulletin of the Atomic Scientists, 56(2):4–5, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
2/4.

Anonymous:2000:Lc

[Ano00k] Anonymous. Letters. Bulletin of the Atomic Scientists, 56(3):3–5, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
3/3.

Anonymous:2000:Ld

[Ano00l] Anonymous. Letters. Bulletin of the Atomic Scientists, 56(4): 3–5, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/4/3.

Anonymous:2000:Le

[Ano00m] Anonymous. Letters. Bulletin of the Atomic Scientists, 56(5): 3, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/5/3.

Anonymous:2000:Lf

[Ano00n] Anonymous. Letters. Bulletin of the Atomic Scientists, 56(6):3, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/6/3.

Anonymous:2000:UNF

[Ano00o] Anonymous. U.S. nuclear forces, 2000. Bulletin of the Atomic Scientists, 56(3):69, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2001:BN

[Ano01a] Anonymous. Book note. Bulletin of the Atomic Scientists, 57 (3):74, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/3/74.

Anonymous:2001:BNF

[Ano01b] Anonymous. Britain nuclear forces, 2001. Bulletin of the Atomic Scientists, 57(6):78, November/December 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2001:Ba

[Ano01c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 57(1):6– 10, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/1/6.

Anonymous:2001:Bb

[Ano01d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 57(2):6– 10, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
2/6.

Anonymous:2001:Bc

[Ano01e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 57 (3):6–13, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/3/6.

Anonymous:2001:Bd

[Ano01f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 57(4): 6–12, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
4/6.

Anonymous:2001:Bf

[Ano01g] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 57(5):6– 12, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/5/6.

Anonymous:2001:Bg

[Ano01h] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 57 (6):6–11, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/6/6.

Anonymous:2001:Be

[Ano01i] Anonymous. In brief. Bulletin of the Atomic Scientists, 57(4):9, July/August 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2001:I

[Ano01j] Anonymous. Index 2001. Bulletin of the Atomic Scientists, 57 (6):73–77, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/6/73.

Anonymous:2001:La

[Ano01k] Anonymous. Letters. Bulletin of the Atomic Scientists, 57(1): ??, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/1/4.

Anonymous:2001:Lb

[Ano01l] Anonymous. Letters. Bulletin of the Atomic Scientists, 57(2):4–5, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
2/4.

Anonymous:2001:Lc

[Ano01m] Anonymous. Letters. Bulletin of the Atomic Scientists, 57(3):4–5, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
3/4.

Anonymous:2001:Ld

[Ano01n] Anonymous. Letters. Bulletin of the Atomic Scientists, 57(4):4–5, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/57/4/4.

Anonymous:2001:Le

[Ano01o] Anonymous. Letters. Bulletin of the Atomic Scientists, 57(5):4– 5, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/5/4.

Anonymous:2001:Lf

[Ano01p] Anonymous. Letters. Bulletin of the Atomic Scientists, 57(6):4–5, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
6/4.

Anonymous:2001:Ua

[Ano01q] Anonymous. Updates. Bulletin of the Atomic Scientists, 57(3): 5, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2001:Ub

[Ano01r] Anonymous. Updates. Bulletin of the Atomic Scientists, 57(4): 5, July/August 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2002:Ba

[Ano02a] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 58(1):6– 13, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/1/6.

Anonymous:2002:Bb

[Ano02b] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 58(2): 10–16, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/2/10.

Anonymous:2002:Bc

[Ano02c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 58 (3):6–12, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/3/6.

Anonymous:2002:Bd

[Ano02d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 58(4): 6–10, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/4/6.

Anonymous:2002:Be

[Ano02e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 58(5):6– 11, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/5/6.

Anonymous:2002:Bf

[Ano02f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 58 (6):6–11, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/6.

Anonymous:2002:EET

[Ano02g] Anonymous. Encounters with Edward Teller. Bulletin of the Atomic Scientists, 58(1):4, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2002:I

[Ano02h] Anonymous. Index 2002. Bulletin of the Atomic Scientists, 58 (6):105–108, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/105.

Anonymous:2002:INF

[Ano02i] Anonymous. India’s nuclear forces, 2002. Bulletin of the Atomic Scientists, 58(2):70, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2002:SMM

[Ano02j] Anonymous. It’s seven minutes to midnight. Bulletin of the Atomic Scientists, 58(2):4–7, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/4.

Anonymous:2002:LP

[Ano02k] Anonymous. Letter from Pyongyang. Bulletin of the Atomic Scientists, 58(4):50–54, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/50.

Anonymous:2002:La

[Ano02l] Anonymous. Letters. Bulletin of the Atomic Scientists, 58(1): ??, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/1/4.

Anonymous:2002:Lb

[Ano02m] Anonymous. Letters. Bulletin of the Atomic Scientists, 58 (2):??, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/2/8.

Anonymous:2002:Lc

[Ano02n] Anonymous. Letters. Bulletin of the Atomic Scientists, 58(3):4–5, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/58/
3/4.

Anonymous:2002:Ld

[Ano02o] Anonymous. Letters. Bulletin of the Atomic Scientists, 58(4): 4–5, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/4/4.

Anonymous:2002:Le

[Ano02p] Anonymous. Letters. Bulletin of the Atomic Scientists, 58(5):4– 5, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/5/4.

Anonymous:2002:Lf

[Ano02q] Anonymous. Letters. Bulletin of the Atomic Scientists, 58(6):4–5, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/58/
6/4.

Anonymous:2003:Ba

[Ano03a] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 59(1):6– 11, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/1/6.

Anonymous:2003:Bb

[Ano03b] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 59(2):6– 11, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
2/6.

Anonymous:2003:Bc

[Ano03c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 59 (3):6–11, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/3/6.

Anonymous:2003:Bd

[Ano03d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 59(4): 6–15, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/4/6.

Anonymous:2003:Be

[Ano03e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 59(5):6– 10, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/5/6.

Anonymous:2003:Bf

[Ano03f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 59(6): 6–11, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/6/6.

Anonymous:2003:DS

[Ano03g] Anonymous. Dubious secrets. Bulletin of the Atomic Scientists, 59(5):36–37, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/5/36.

Anonymous:2003:GSU

[Ano03h] Anonymous. Gimme shelter: Underground America. Bulletin of the Atomic Scientists, 59(3):38–39, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/3/38.

Anonymous:2003:HSH

[Ano03i] Anonymous. Hanford Site historic district: history of the plutonium production facilities, 1943–1990. Battelle Press, Columbus, OH, USA, 2003. ISBN 1-57477-133-7 (paperback). ???? pp. LCCN TK9025 .H36 2003. URL http://www.h-net.org/

review/hrev-a0d7n2-aa.

Anonymous:2003:I

[Ano03j] Anonymous. Index 2003. Bulletin of the Atomic Scientists, 59(6): 81–84, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/6/81.

Anonymous:2003:La

[Ano03k] Anonymous. Letters. Bulletin of the Atomic Scientists, 59(1): 4–5, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/1/4.

Anonymous:2003:Lb

[Ano03l] Anonymous. Letters. Bulletin of the Atomic Scientists, 59(2):4–5, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
2/4.

Anonymous:2003:Lc

[Ano03m] Anonymous. Letters. Bulletin of the Atomic Scientists, 59(3):??, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
3/4.

Anonymous:2003:Ld

[Ano03n] Anonymous. Letters. Bulletin of the Atomic Scientists, 59(4): 4–5, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/4/4.

Anonymous:2003:Le

[Ano03o] Anonymous. Letters. Bulletin of the Atomic Scientists, 59(5):??, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/5/4.

Anonymous:2003:Lf

[Ano03p] Anonymous. Letters. Bulletin of the Atomic Scientists, 59(6):??, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/6/4.

Anonymous:2004:Ba

[Ano04a] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 60 (1):6–14, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/1/6.

Anonymous:2004:Bb

[Ano04b] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 60(2): 6–11, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/2/6.

Anonymous:2004:Bc

[Ano04c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 60 (3):6–12, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/3/6.

Anonymous:2004:Bd

[Ano04d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 60(4): 6–11, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/4/6.

Anonymous:2004:Be

[Ano04e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 60(5):6– 10, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/5/6.

Anonymous:2004:Bg

[Ano04f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 60(6): 6–11, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/6/6.

Anonymous:2004:Bf

[Ano04g] Anonymous. In brief. Bulletin of the Atomic Scientists, 60(5):9– 10, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2004:I

[Ano04h] Anonymous. Index 2004. Bulletin of the Atomic Scientists, 60(6): 81–84, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/6/81.

Anonymous:2004:La

[Ano04i] Anonymous. Letters. Bulletin of the Atomic Scientists, 60(1):??, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/60/
1/4.

Anonymous:2004:Lb

[Ano04j] Anonymous. Letters. Bulletin of the Atomic Scientists, 60(2): 4–5, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/2/4.

Anonymous:2004:Lc

[Ano04k] Anonymous. Letters. Bulletin of the Atomic Scientists, 60(3):??, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/60/
3/4.

Anonymous:2004:Ld

[Ano04l] Anonymous. Letters. Bulletin of the Atomic Scientists, 60 (4):??, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/4/4.

Anonymous:2004:Le

[Ano04m] Anonymous. Letters. Bulletin of the Atomic Scientists, 60(5):4– 5, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/5/4.

Anonymous:2004:Lf

[Ano04n] Anonymous. Letters. Bulletin of the Atomic Scientists, 60(6):5, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/6/5.

Anonymous:2005:W

[Ano05a] Anonymous. And the winner is. Bulletin of the Atomic Scientists, 61(6):6, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2005:BRB

[Ano05b] Anonymous. Book review: Edward Teller: the Real Dr. Strangelove, by Peter Goodchild. Bulletin of the Atomic Scientists, 61(2):69, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:/

/www.highbeam.com/doc/1G1-129710745.html. See [Goo04a,

Goo04b].

Anonymous:2005:Ba

[Ano05c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 61(1):6– 11, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/1/6.

Anonymous:2005:Bb

[Ano05d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 61(2): 6–11, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/2/6.

Anonymous:2005:Bc

[Ano05e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 61 (3):8–13, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/3/8.

Anonymous:2005:Bd

[Ano05f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 61(4): 8–13, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/4/8.

Anonymous:2005:Be

[Ano05g] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 61(5):8– 13, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/5/8.

Anonymous:2005:Bf

[Ano05h] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 61(6): 8–13, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/6/8.

Anonymous:2005:I

[Ano05i] Anonymous. Index 2005. Bulletin of the Atomic Scientists, 61(6): 71–74, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/6/71.

Anonymous:2005:La

[Ano05j] Anonymous. Letters. Bulletin of the Atomic Scientists, 61(1): 5, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/1/5.

Anonymous:2005:Lb

[Ano05k] Anonymous. Letters. Bulletin of the Atomic Scientists, 61 (2):??, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/2/5.

Anonymous:2005:Lc

[Ano05l] Anonymous. Letters. Bulletin of the Atomic Scientists, 61(3):7, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
3/7.

Anonymous:2005:Ld

[Ano05m] Anonymous. Letters. Bulletin of the Atomic Scientists, 61 (4):??, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/4/5.

Anonymous:2005:Le

[Ano05n] Anonymous. Letters. Bulletin of the Atomic Scientists, 61(5):5– 7, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/5/5.

Anonymous:2005:Lf

[Ano05o] Anonymous. Letters. Bulletin of the Atomic Scientists, 61(6):5– 7, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/6/5.

Anonymous:2005:P

[Ano05p] Anonymous. Passings. Bulletin of the Atomic Scientists, 61(6): 14–15, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/6/14.

Anonymous:2006:Ba

[Ano06a] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 62(1):8– 13, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/1/8.

Anonymous:2006:Bb

[Ano06b] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 62(2): 8–13, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/2/8.

Anonymous:2006:Bc

[Ano06c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 62 (3):9–13, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/3/9.

Anonymous:2006:Bd

[Ano06d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 62(4): 9–13, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/4/9.

Anonymous:2006:Be

[Ano06e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 62(5):9– 13, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/5/9.

Anonymous:2006:Bf

[Ano06f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 62(6): 9–13, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/6/9.

Anonymous:2006:C

[Ano06g] Anonymous. The clock. Bulletin of the Atomic Scientists, 62(3): 5, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2006:Bg

[Ano06h] Anonymous. In brief. Bulletin of the Atomic Scientists, 62(6): 10–12, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2006:I

[Ano06i] Anonymous. Index 2006. Bulletin of the Atomic Scientists, 62(6): 59–61, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/6/59.

Anonymous:2006:La

[Ano06j] Anonymous. Letters. Bulletin of the Atomic Scientists, 62(1): 5–7, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/1/5.

Anonymous:2006:Lb

[Ano06k] Anonymous. Letters. Bulletin of the Atomic Scientists, 62(2): 5–7, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/2/5.

Anonymous:2006:Lc

[Ano06l] Anonymous. Letters. Bulletin of the Atomic Scientists, 62(3):5–8, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/62/
3/5.

Anonymous:2006:Ld

[Ano06m] Anonymous. Letters. Bulletin of the Atomic Scientists, 62(4): 5–8, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/4/5.

Anonymous:2006:Le

[Ano06n] Anonymous. Letters. Bulletin of the Atomic Scientists, 62(5):5– 8, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/5/5.

Anonymous:2006:Lf

[Ano06o] Anonymous. Letters. Bulletin of the Atomic Scientists, 62(6):5– 8, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/6/5.

Anonymous:2006:P

[Ano06p] Anonymous. Passings. Bulletin of the Atomic Scientists, 62(6): 14–16, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/6/14.

Anonymous:2006:VE

[Ano06q] Anonymous. A violent escalation. Bulletin of the Atomic Scientists, 62(6):6, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2007:Ba

[Ano07a] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 63(1):9– 12, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/1/9.

Anonymous:2007:Bb

[Ano07b] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 63(2): 11–14, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/2/11.

Anonymous:2007:Bc

[Ano07c] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 63 (3):11–17, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/3/11.

Anonymous:2007:Bd

[Ano07d] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 63 (4):11–15, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/4/11.

Anonymous:2007:Be

[Ano07e] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 63 (5):9–13, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/5/9.

Anonymous:2007:Bf

[Ano07f] Anonymous. Bulletins. Bulletin of the Atomic Scientists, 63(6): 9–15, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/6/9.

Anonymous:2007:CS

[Ano07g] Anonymous. The clock since 1947. Bulletin of the Atomic Scientists, 63(1):68–69, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2007:HHS

[Ano07h] Anonymous. Hunting for hot spots. Bulletin of the Atomic Scientists, 63(2):22–23, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/2/22.

Anonymous:2007:IC

[Ano07i] Anonymous. The Indian claim. Bulletin of the Atomic Scientists, 63(1):9, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2007:La

[Ano07j] Anonymous. Letters. Bulletin of the Atomic Scientists, 63(1): 5–8, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/1/5.

Anonymous:2007:Lb

[Ano07k] Anonymous. Letters. Bulletin of the Atomic Scientists, 63(2): 5–9, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/2/5.

Anonymous:2007:Lc

[Ano07l] Anonymous. Letters. Bulletin of the Atomic Scientists, 63(3):5–9, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
3/5.

Anonymous:2007:Ld

[Ano07m] Anonymous. Letters. Bulletin of the Atomic Scientists, 63(4):5–9, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/63/4/5.

Anonymous:2007:Le

[Ano07n] Anonymous. Letters. Bulletin of the Atomic Scientists, 63 (5):5–8, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/5/5.

Anonymous:2007:Lf

[Ano07o] Anonymous. Letters. Bulletin of the Atomic Scientists, 63(6):5– 8, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/6/5.

Anonymous:2007:OHH

[Ano07p] Anonymous. Old habits die hard. Bulletin of the Atomic Scientists, 63(1):38, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2007:PC

[Ano07q] Anonymous. The Pakistani claim. Bulletin of the Atomic Scientists, 63(1):9, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2007:WB

[Ano07r] Anonymous. What is biosecurity? Bulletin of the Atomic Scientists, 63(2):18–19, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/2/18.

Anonymous:2008:WTOa

[Ano08a] Anonymous. At www.thebulletin.org. Bulletin of the Atomic Scientists, 64(1):9, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/64/1/9.

Anonymous:2008:WTOb

[Ano08b] Anonymous. At www.thebulletin.org. Bulletin of the Atomic Scientists, 64(2):8, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/64/2/8.

Anonymous:2008:La

[Ano08c] Anonymous. Letters. Bulletin of the Atomic Scientists, 64(1): 4–5, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/1/4.

Anonymous:2008:Lb

[Ano08d] Anonymous. Letters. Bulletin of the Atomic Scientists, 64(2):4–6, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/64/
2/4.

Anonymous:2008:Lc

[Ano08e] Anonymous. Letters. Bulletin of the Atomic Scientists, 64(3): 6–7, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/3/6.

Anonymous:2008:Ld

[Ano08f] Anonymous. Letters. Bulletin of the Atomic Scientists, 64 (4):4–6, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/4/4.

Anonymous:2008:Le

[Ano08g] Anonymous. Letters. Bulletin of the Atomic Scientists, 64(5):4– 5, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/5/4.

Anonymous:2008:PO

[Ano08h] Anonymous. Peak oil. Bulletin of the Atomic Scientists, 64(5):??, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/5/34.

Anonymous:2009:IMP

[Ano09] Anonymous. Interview: Michael Polsky. Bulletin of the Atomic Scientists, 65(6):72–77, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/65/6/72.

Anonymous:2010:BRS

[Ano10a] Anonymous. Book review: Subrata Ghoshroy and Götz Neuneck (eds), South Asia at a Crossroads: Conflict or Cooperation in the Age of Nuclear Weapons, Missile Defense, and Space Rivalries. Baden-Baden: Nomos, 2010. 330 pp. $86.00. ISBN 978-3-8329-5398-0. Bulletin of the Atomic Scientists, 66(6):130–133, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/6/130.

Anonymous:2010:GAC

[Ano10b] Anonymous. Graham T. Allison: The congenital optimist. Bulletin of the Atomic Scientists, 66(5):1–8, September/October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/5/1.

Anonymous:2010:MMS

[Ano10c] Anonymous. Michael E. Mann: a scientist in the crosshairs of climate-change denial. Bulletin of the Atomic Scientists, 66(6): 1–7, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/66/6/1.

Anonymous:2011:Ea

[Ano11a] Anonymous. Erratum. Bulletin of the Atomic Scientists, 67(1): 76, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/1/76.

Anonymous:2011:Eb

[Ano11b] Anonymous. Erratum. Bulletin of the Atomic Scientists, 67(5): 98, September/October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/5/98.

Anonymous:2011:GSC

[Ano11c] Anonymous. Gus Speth: Communicating environmental risks in an age of disinformation. Bulletin of the Atomic Scientists, 67 (4):1–7, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/4/1.

Anonymous:2011:OHR

[Ano11d] Anonymous. Olli Heinonen: Reporting from the front lines of nuclear proliferation. Bulletin of the Atomic Scientists, 67(2): 1–9, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/2/1.

Anonymous:2011:PRE

[Ano11e] Anonymous. Paul R. Ehrlich: Seeking environmental solutions in the social sciences. Bulletin of the Atomic Scientists, 67(5):1– 8, September/October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/5/1.

Anonymous:2011:RDT

[Ano11f] Anonymous. Ronald Deibert: Tracking the emerging arms race in cyberspace. Bulletin of the Atomic Scientists, 67(1):1–8, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/1/1.

Anonymous:2011:RGG

[Ano11g] Anonymous. Rose Gottemoeller: Getting to yes. Bulletin of the Atomic Scientists, 67(6):1–8, November/December 2011. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/67/6/1.

Anonymous:2011:SOP

[Ano11h] Anonymous. Stanford Ovshinsky: Pursuing solar electricity at a cost equal to or lower than that of coal electricity. Bulletin of the Atomic Scientists, 67(3):1–7, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/67/3/1.

Anonymous:2012:BMA

[Ano12a] Anonymous. Bill McKibben: Actions speak louder than words. Bulletin of the Atomic Scientists, 68(2):1–8, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/2/1.

Anonymous:2012:BRS

[Ano12b] Anonymous. Burton Richter: Seeing energy in three dimensions. Bulletin of the Atomic Scientists, 68(4):1–8, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/4/1.

Anonymous:2012:C

[Ano12c] Anonymous. Corrigendum. Bulletin of the Atomic Scientists, 68 (4):102, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/4/102.

Anonymous:2012:LGR

[Ano12d] Anonymous. Laurie Garrett: Reporting on biosecurity from America to Zaire. Bulletin of the Atomic Scientists, 68(1):1– 9, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/1/1.

Anonymous:2012:OZA

[Ano12e] Anonymous. Ozzie Zehner: Alternatives to alternative energy. Bulletin of the Atomic Scientists, 68(5):1–7, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/5/1.

Anonymous:2012:THD

[Ano12f] Anonymous. Thomas Homer-Dixon: Exploring the climate “mindscape”. Bulletin of the Atomic Scientists, 68(3):1–9, May/ June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/68/3/1.

Anonymous:2012:WNN

[Ano12g] Anonymous. What’s next for the NRC? A conversation with Allison Macfarlane. Bulletin of the Atomic Scientists, 68(6):1–5, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/6/1.

Anonymous:2013:GPS

[Ano13a] Anonymous. George P. Shultz: a Cold Warrior on a warming planet. Bulletin of the Atomic Scientists, 69(1):1–8, January/ February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/69/
1/1.

Anonymous:2013:GPO

[Ano13b] Anonymous. George Poste: One step ahead. Bulletin of the Atomic Scientists, 69(6):1–10, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/6/1.

Anonymous:2013:ISH

[Ano13c] Anonymous. Interview: Siegfried S. Hecker: The story of Plutonium Mountain. Bulletin of the Atomic Scientists, 69 (5):1–7, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/5/1.

Anonymous:2013:JHT

[Ano13d] Anonymous. James Hansen: Taking heat for decades. Bulletin of the Atomic Scientists, 69(4):1–8, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/69/4/1.

Anonymous:2013:KHP

[Ano13e] Anonymous. Katharine Hayhoe: Preaching climate to the unconverted. Bulletin of the Atomic Scientists, 69(3):1–9, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/3/1.

Anonymous:2013:OMH

[Ano13f] Anonymous. Osamu Motojima: Harnessing the 93-million-mile dream. Bulletin of the Atomic Scientists, 69(2):1–8, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/2/1.

Anonymous:2014:CHM

[Ano14a] Anonymous. Corrigendum: Hans M. Kristensen and Robert S. Norris. (2014) Russian nuclear forces, 2014. Bulletin of the Atomic Scientists 70(2): 75–85. Bulletin of the Atomic Scientists, 70(4):108, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [KN14c].

Anonymous:2014:EKC

[Ano14b] Anonymous. Elizabeth Kolbert: Covering the hot topic of climate change by going to the ends of the Earth. Bulletin of the Atomic Scientists, 70(4):1–9, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2014:FCY

[Ano14c] Anonymous. Frances Crowe, 95-year-old antinuclear activist. Bulletin of the Atomic Scientists, 70(6):1–9, November 2014. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2014:ESU

[Ano14d] Anonymous. Interview: Eric Schlosser: Uncovering nuclear weapons history from the ground up. Bulletin of the Atomic Scientists, 70(2):1–8, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2014:JBC

[Ano14e] Anonymous. Jerry Brown: Climate change policy in California — and beyond. Bulletin of the Atomic Scientists, 70(5):1–7, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2014:LBP

[Ano14f] Anonymous. Lester Brown: Plowing the way in environmental research. Bulletin of the Atomic Scientists, 70(1):1–8, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/70/1/1.
full.pdf+html.

Anonymous:2014:TWW

[Ano14g] Anonymous. Tom Wigley: Why nuclear power may be the only way to avoid geoengineering. Bulletin of the Atomic Scientists, 70 (3):10–16, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:CCC

[Ano15a] Anonymous. Cloud control: Climatologist Alan Robock on the effects of geoengineering and nuclear war. Bulletin of the Atomic Scientists, 71(3):1–7, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:FHS

[Ano15b] Anonymous. Frank von Hippel, scientist in the public interest. Bulletin of the Atomic Scientists, 71(1):1–9, January 2015. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:GCW

[Ano15c] Anonymous. Gwyneth Cravens: Why nuclear power should play a greater role in the response to climate change. Bulletin of the Atomic Scientists, 71(6):45–51, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:HJS

[Ano15d] Anonymous. Henry Jacqz: a student of climate change. Bulletin of the Atomic Scientists, 71(2):1–8, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:JRW

[Ano15e] Anonymous. Joe Romm: Why nuclear power will not be the whole solution to climate change. Bulletin of the Atomic Scientists, 71(6):52–58, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:MTM

[Ano15f] Anonymous. Mitchie Takeuchi and Miyako Taguchi: Secondgeneration survivors of the atomic bomb. Bulletin of the Atomic Scientists, 71(4):1–9, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2015:RNP

[Ano15g] Anonymous. A religious nature: Philosopher Seyyed Hossein Nasr on Islam and the environment. Bulletin of the Atomic Scientists, 71(5):13–18, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/13.full.pdf+html.

Anonymous:2016:EB

[Ano16a] Anonymous. Editorial board. Bulletin of the Atomic Scientists, 72(6):ebi, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2016:EN

[Ano16b] Anonymous. Editor’s note. Bulletin of the Atomic Scientists, 72 (1):1, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2016:E

[Ano16c] Anonymous. Erratum. Bulletin of the Atomic Scientists, 72(4): 222–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2017:A

[Ano17a] Anonymous. Addendum. Bulletin of the Atomic Scientists, 73(2):143, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1290937.

Anonymous:2017:C

[Ano17b] Anonymous. Corrigendum. Bulletin of the Atomic Scientists, 73(2):144, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1290934.

Anonymous:2017:EB

[Ano17c] Anonymous. Editorial board. Bulletin of the Atomic Scientists, 73(6):??, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1395597.

Anonymous:2018:CUG

[Ano18] Anonymous. Corrigendum: [US ground-based midcourse missile defense: Expensive and unreliable]. Bulletin of the Atomic Scientists, 74(5):359, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2018.1495860. See [Gre18].

Anonymous:2019:Ca

[Ano19a] Anonymous. Correction. Bulletin of the Atomic Scientists, 75 (4):210–212, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1612205.

Anonymous:2019:Cb

[Ano19b] Anonymous. Correction. Bulletin of the Atomic Scientists, 75(4):213, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1627773.

Anonymous:2019:Cc

[Ano19c] Anonymous. Correction. Bulletin of the Atomic Scientists, 75(5):262, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1640502.

Anonymous:2020:C

[Ano20] Anonymous. Correction. Bulletin of the Atomic Scientists, 76(2):118, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1725323.

Anonymous:2022:Ca

[Ano22a] Anonymous. Correction. Bulletin of the Atomic Scientists, 78(5):295, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2103283.

Anonymous:2022:Cb

[Ano22b] Anonymous. Correction. Bulletin of the Atomic Scientists, 79 (1):53–54, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2147705.

Anonymous:2022:WWH

[Ano22c] Anonymous. William “Willy” Higinbotham: Physicist, Los Alamos, NM. Atomic Heritage Foundation Web site, 2022. URL

https://ahf.nuclearmuseum.org/ahf/profile/william-willy-
higinbotham/.

Anthony:1950:PNC

[Ant50] S. A. Anthony, Jr. Preparing Now for civil defense responsibility. Bulletin of the Atomic Scientists, 6(8–9):271–272, August/ September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Anthrop:1969:ENP

[Ant69] Donald F. Anthrop. Environmental noise pollution: A new threat to sanity. Bulletin of the Atomic Scientists, 25(5):11–16, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anthrop:1970:ESE

[Ant70] Donald F. Anthrop. Environmental side effects of energy production. Bulletin of the Atomic Scientists, 26(8):39–41, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anthrop:1974:NLT

[Ant74] Donald F. Anthrop. The need for a long-term policy: Nixon Administration is headed the wrong way. Bulletin of the Atomic Scientists, 30(5):33–38, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Antal:1976:TPP

[Ant76] M. J. Antal, Jr. Tower power: producing fuels from solar energy. Bulletin of the Atomic Scientists, 32(5):59–62, May 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Whi76a].

Anthrop:1978:CEP

[Ant78] Donald F. Anthrop. The Carter energy plan and the American West. Bulletin of the Atomic Scientists, 34(1):27–33, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1995:NFY

[AO95a] David Albright and Kevin O’Neill. The next fifty years. Bulletin of the Atomic Scientists, 51(1):24–25, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1995:NJR

[AO95b] David Albright and Kevin O’Neill. Nonproliferation: Jury-rigged, but working. Bulletin of the Atomic Scientists, 51(1):20–26, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1995:PN

[AO95c] David Albright and Kevin O’Neill. The price of nonproliferation. Bulletin of the Atomic Scientists, 51(1):27–29, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kim95].

Albright:1998:IRS

[AO98] David Albright and Kevin O’Neill. Iraq: Resettle the scientists. Bulletin of the Atomic Scientists, 54(1):17, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1980:PTD

[AP80] Nancy E. Abrams and Joel R. Primack. The public and technological decisions. Bulletin of the Atomic Scientists, 36(6):44–48, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cop80].

Albert:1988:MWC

[AP88] David Albert and Christopher Paine. Making warheads: A case against producing nuclear material. Bulletin of the Atomic Scientists, 44(1):46–49, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Appleton:1946:SGI

[App46] Sir Edward V. Appleton. Science, government and industry in Britain. Bulletin of the Atomic Scientists, 2(11–12):2–5, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Appleton:1953:SOS

[App53] Sir Edward Appleton. Science for its own sake. Bulletin of the Atomic Scientists, 9(10):354–358, 390, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the Presidential Address in Liverpool to the British Association in September 1953.

Anichkina:2017:FUR

[APR17] Tatiana Anichkina, Anna Péczeli, and Nickolas Roth. The future of US–Russian nuclear deterrence and arms control. Bulletin of the Atomic Scientists, 73(4):271–278, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1338046.

Apter:1961:NNS

[Apt61] David E. Apter. New nations and the scientific revolution. Bulletin of the Atomic Scientists, 17(2):60–64, February 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aptheker:1964:LET

[Apt64] Herbert Aptheker. Letter to the Editor: To travel freely. Bulletin of the Atomic Scientists, 20(3):25, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1986:DMR

[APvH86] David Albright, Christopher Paine, and Frank von Hippel. The danger of military reactors. Bulletin of the Atomic Scientists, 42(8):44–48, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1980:CMA

[AR80] Robert Alvarez and Joseph Rotblat. Commentary: Mancuso affair. Bulletin of the Atomic Scientists, 36(2):61–63, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot79a].

Alexander:1994:LKM

[ARA94a] John B. Alexander, Gene Rochlin, and Steven Aftergood. Letters: Kill machines, not men. Bulletin of the Atomic Scientists, 50(6): 3, 56, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Aft94b].

Arana:1994:SC

[Ara94b] Ana Arana. The smiling chameleon. Bulletin of the Atomic Scientists, 50(6):36–39, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbatov:1977:DCW

[Arb77] Georgi Arbatov. The dangers of the Cold War. Bulletin of the Atomic Scientists, 33(3):33–40, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Lic77, Haa77].

Arbatov:1982:VM

[Arb82] Georgi Arbatov. View from Moscow. Bulletin of the Atomic Scientists, 38(9):8–9, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbess:1985:SWO

[Arb85] Daniel Arbess. Star Wars and outer space law. Bulletin of the Atomic Scientists, 41(9):19–22, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbatov:1991:SMM

[Arb91a] Alexei G. Arbatov. START: Monitoring mobiles. Bulletin of the Atomic Scientists, 47(9):38, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbatov:1991:SWC

[Arb91b] Alexei G. Arbatov. START: We could have done better. Bulletin of the Atomic Scientists, 47(9):36–40, 47, November 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbatov:1993:SIR

[Arb93] Alexei Arbatov. START-II, Red ink, and Boris Yeltsin. Bulletin of the Atomic Scientists, 49(3):16–21, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbatov:2016:SNA

[Arb16] Alexei Arbatov. Saving nuclear arms control. Bulletin of the Atomic Scientists, 72(3):165–170, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arbatov:2018:VRM

[Arb18] Alexey Arbatov. The vicissitudes of Russian missile defense. Bulletin of the Atomic Scientists, 74(4):227–237, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1486595.

Arbatov:2019:ACO

[Arb19] Alexey Arbatov. Arms control in outer space: The Russian angle, and a possible way forward. Bulletin of the Atomic Scientists, 75(4):151–161, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1628475.

York:1954:UMS

[Arc54] Archbishop of York. The use and misuse of science. Bulletin of the Atomic Scientists, 10(5):155, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Archer:1986:UMW

[Arc86] Dane Archer. University management of weapons labs? No. Bulletin of the Atomic Scientists, 42(1):41–44, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See opposing view [Kah86].

Ardaev:1993:A

[Ard93a] Vladimir Ardaev. AWOL. Bulletin of the Atomic Scientists, 49 (8):36, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ardaev:1993:BEW

[Ard93b] Vladimir Ardaev. Bridging east and west. Bulletin of the Atomic Scientists, 49(8):23–29, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ardaev:1993:DBM

[Ard93c] Vladimir Ardaev. Don’t blame Moscow. Bulletin of the Atomic Scientists, 49(8):52, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1983:WNS

[Ark83a] William M. Arkin. From Washington: Nuclear security: the enemy may be us. Bulletin of the Atomic Scientists, 39(9):4–5, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1983:WNW

[Ark83b] William M. Arkin. From Washington: Nuclear weapons at sea. Bulletin of the Atomic Scientists, 39(8):6–7, October 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1983:WNB

[Ark83c] William M. Arkin. From Washington: The nuclear balancing act in the Pacific. Bulletin of the Atomic Scientists, 39(10):9– 10, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1983:PIU

[Ark83d] William M. Arkin. Pershing II and U.S. nuclear strategy. Bulletin of the Atomic Scientists, 39(6):12–13, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1983:WS

[Ark83e] William M. Arkin. Why SIOP-6? Bulletin of the Atomic Scientists, 39(4):9–10, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:WAM

[Ark84a] William M. Arkin. From Washington: Arsenals move north. Bulletin of the Atomic Scientists, 40(3):5–6, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:WFF

[Ark84b] William M. Arkin. From Washington: Flying in the face of arms control. Bulletin of the Atomic Scientists, 40(2):5–6, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:WGS

[Ark84c] William M. Arkin. From Washington: Going with small ICBMs. Bulletin of the Atomic Scientists, 40(5):7–8, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:WSP

[Ark84d] William M. Arkin. From Washington: SDI — pie in the sky? Bulletin of the Atomic Scientists, 40(4):9–10, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:PTO

[Ark84e] William M. Arkin. Perspective: Tomahawk: ominous new development. Bulletin of the Atomic Scientists, 40(8):3–4, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1984:SHT

[Ark84f] William M. Arkin. Sleight of hand with Trident II. Bulletin of the Atomic Scientists, 40(10):5–6, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cas85b].

Arkin:1985:CBD

[Ark85a] William M. Arkin. Conventional buildup — a deliberate delusion. Bulletin of the Atomic Scientists, 41(2):8–9, February 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRB

[Ark85b] William M. Arkin. Washington report: Beltway bandits. Bulletin of the Atomic Scientists, 41(8):5–6, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRC

[Ark85c] William M. Arkin. Washington report: Confirm or deny. Bulletin of the Atomic Scientists, 41(11):4–5, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRN

[Ark85d] William M. Arkin. Washington report: Nuclear backpacks. Bulletin of the Atomic Scientists, 41(4):4–5, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRPb

[Ark85e] William M. Arkin. Washington report: Playing chicken in Turkey. Bulletin of the Atomic Scientists, 41(9):4–5, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRPa

[Ark85f] William M. Arkin. Washington report: Preparing for World War IV. Bulletin of the Atomic Scientists, 41(5):6–7, May 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRPc

[Ark85g] William M. Arkin. Washington report: Provocations at sea. Bulletin of the Atomic Scientists, 41(10):6–7, November 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1985:WRD

[Ark85h] William M. Arkin. Washington report: The drift toward first strike. Bulletin of the Atomic Scientists, 41(1):5–6, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho85a].

Arkin:1985:WRW

[Ark85i] William M. Arkin. Washington report: Waging secrecy. Bulletin of the Atomic Scientists, 41(3):5–6, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRG

[Ark86a] William M. Arkin. Washington report: A global role for NATO. Bulletin of the Atomic Scientists, 42(1):4–5, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRM

[Ark86b] William M. Arkin. Washington report: At the movies. Bulletin of the Atomic Scientists, 42(5):6–7, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRC

[Ark86c] William M. Arkin. Washington report: Canada — too close for comfort. Bulletin of the Atomic Scientists, 42(3):4–5, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRF

[Ark86d] William M. Arkin. Washington report: Fewer warheads in Europe. Bulletin of the Atomic Scientists, 42(7):4–5, August/ September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRNb

[Ark86e] William M. Arkin. Washington report: No use for no first use. Bulletin of the Atomic Scientists, 42(9):4–5, November 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRNc

[Ark86f] William M. Arkin. Washington report: Nuclear war in triplicate. Bulletin of the Atomic Scientists, 42(10):6–7, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRD

[Ark86g] William M. Arkin. Washington report: Of drugs and Star Wars. Bulletin of the Atomic Scientists, 42(2):4–5, February 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRT

[Ark86h] William M. Arkin. Washington report: Test ban fever. Bulletin of the Atomic Scientists, 42(8):4–5, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1986:WRNa

[Ark86i] William M. Arkin. Washington report: The new mix of defense and deterrence. Bulletin of the Atomic Scientists, 42(6): 4–5, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1987:ARSa

[Ark87a] William M. Arkin. Arms race at sea: Navy autonomy thwarts arms control. Bulletin of the Atomic Scientists, 43(7):14–18, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1987:ARSb

[Ark87b] William M. Arkin. Arms race at sea: The Department of the Navy as subsidiary. Bulletin of the Atomic Scientists, 43(7):17, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1987:WRGa

[Ark87c] William M. Arkin. Washington report: General Nutting and the invaders. Bulletin of the Atomic Scientists, 43(1):6–7, January/ February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1987:WRGb

[Ark87d] William M. Arkin. Washington report: Greece’s balancing act. Bulletin of the Atomic Scientists, 43(2):11–12, March 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1987:WRH

[Ark87e] William M. Arkin. Washington report: Happy birthday, flexible response. Bulletin of the Atomic Scientists, 43(10):5–6, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1987:WRL

[Ark87f] William M. Arkin. Washington report: Long on data, short on intelligence. Bulletin of the Atomic Scientists, 43(5):5–6, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zim87].

Arkin:1987:WRP

[Ark87g] William M. Arkin. Washington report: Pentagon banking on Castro phobia. Bulletin of the Atomic Scientists, 43(4):4–5, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wie87].

Arkin:1987:WRR

[Ark87h] William M. Arkin. Washington report: Red herring in the Pacific. Bulletin of the Atomic Scientists, 43(3):6–7, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1988:SBC

[Ark88a] William M. Arkin. Spyship Caron: a chronology. Bulletin of the Atomic Scientists, 44(4):6, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1988:WRS

[Ark88b] William M. Arkin. Washington report: Spying in the Black Sea. Bulletin of the Atomic Scientists, 44(4):5–6, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1989:BKO

[Ark89a] William M. Arkin. Bulletin: Keep our secrets and theirs, too. Bulletin of the Atomic Scientists, 45(9):3, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1989:GTW

[Ark89b] William M. Arkin. Gorbachev talks but who listens? Bulletin of the Atomic Scientists, 45(2):5–6, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1989:NWB

[Ark89c] William M. Arkin. Nuclear weapons: The buildup that wasn’t. Bulletin of the Atomic Scientists, 45(1):6–10, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1989:WRS

[Ark89d] William M. Arkin. Washington report: Stealth cruise sneaks into Canada. Bulletin of the Atomic Scientists, 45(4):6–7, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ver89, Hau89].

Arkin:1991:BNB

[Ark91] William M. Arkin. Bulletin: Nuclear blues. Bulletin of the Atomic Scientists, 47(5):5–7, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1992:PUC

[Ark92] William M. Arkin. Perspective: Unindicted co-conspirators. Bulletin of the Atomic Scientists, 48(6):12–13, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1993:NJL

[Ark93a] William M. Arkin. Nuclear junkies: Those lovable little bombs. Bulletin of the Atomic Scientists, 49(6):22–27, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1993:RGW

[Ark93b] William M. Arkin. Report: Gulf War: The desert glows — with propaganda. Bulletin of the Atomic Scientists, 49(4):11–12, 46, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WA93].

Arkin:1994:LWB

[Ark94a] William M. Arkin. The last word: Bad posture. Bulletin of the Atomic Scientists, 50(4):64, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1994:LWMb

[Ark94b] William M. Arkin. The last word: Minibus, son of Mini-nuke. Bulletin of the Atomic Scientists, 50(6):61, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ros94, BT94].

Arkin:1994:LWS

[Ark94c] William M. Arkin. The last word: Success phobia. Bulletin of the Atomic Scientists, 50(5):64, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1994:LWMa

[Ark94d] William M. Arkin. The last word: The 30-minute world. Bulletin of the Atomic Scientists, 50(3):64, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1994:SSF

[Ark94e] William M. Arkin. The sky-is-still-falling profession. Bulletin of the Atomic Scientists, 50(2):64, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1995:LWF

[Ark95a] William M. Arkin. The last word: A fine garble. Bulletin of the Atomic Scientists, 51(3):72, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1995:LWTa

[Ark95b] William M. Arkin. The last word: A tale of two Franks. Bulletin of the Atomic Scientists, 51(2):80, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1995:LWI

[Ark95c] William M. Arkin. The last word: Iran in the cross-hairs. Bulletin of the Atomic Scientists, 51(4):80, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1995:LWP

[Ark95d] William M. Arkin. The last word: Pentagon sees the light: Defense changes its policy on blinding lasers — but still continues to order them. Bulletin of the Atomic Scientists, 51(6):76, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1995:LWTb

[Ark95e] William M. Arkin. The last word: Tales from nuclear storyland: a reader’s guide to the nuclear news. Bulletin of the Atomic Scientists, 51(5):64, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1996:LWNa

[Ark96a] William M. Arkin. The last word: New, and stupid. Bulletin of the Atomic Scientists, 52(1):64, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1996:LWNb

[Ark96b] William M. Arkin. The last word: Nuclear excuses. Bulletin of the Atomic Scientists, 52(3):64, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1996:LWNc

[Ark96c] William M. Arkin. The last word: Nuking Libya. Bulletin of the Atomic Scientists, 52(4):64, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1996:LWC

[Ark96d] William M. Arkin. The last word: The clock strikes out. Bulletin of the Atomic Scientists, 52(2):64, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MT96].

Arkin:1996:LWS

[Ark96e] William M. Arkin. The last word: The six-hundred million dollar mouse. Bulletin of the Atomic Scientists, 52(6):68, November/ December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1996:NE

[Ark96f] William M. Arkin. Nuclear excuses. Bulletin of the Atomic Scientists, 52(3):64, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1996:SWU

[Ark96g] William M. Arkin. The special-weapons umbrella. Bulletin of the Atomic Scientists, 52(5):64, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:BMF

[Ark97a] William M. Arkin. The bomb has many friends. Bulletin of the Atomic Scientists, 53(2):37–39, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:LWBb

[Ark97b] William M. Arkin. The last word: Blame the Internet. Bulletin of the Atomic Scientists, 53(3):65, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:LWBa

[Ark97c] William M. Arkin. The last word: Bring on the radiation. Bulletin of the Atomic Scientists, 53(1):72, January/February 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:LWL

[Ark97d] William M. Arkin. The last word: Leave no stone unburned. Bulletin of the Atomic Scientists, 53(4):64, July/August 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:LWS

[Ark97e] William M. Arkin. The last word: Still glowing. Bulletin of the Atomic Scientists, 53(5):64, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mes97, CA98a].

Arkin:1997:LWBc

[Ark97f] William M. Arkin. The last word: The beat goes on. Bulletin of the Atomic Scientists, 53(6):72, November/December 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:LWT

[Ark97g] William M. Arkin. The last word: The three no’s. Bulletin of the Atomic Scientists, 53(2):64, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1997:WN

[Ark97h] William M. Arkin. What’s “new”? Bulletin of the Atomic Scientists, 53(6):22–27, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:LWC

[Ark98a] William M. Arkin. The last word: Crossed t’s, dotted i’s. Bulletin of the Atomic Scientists, 54(2):72, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:LWFb

[Ark98b] William M. Arkin. The last word: Flying solo. Bulletin of the Atomic Scientists, 54(6):72, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:LWFa

[Ark98c] William M. Arkin. The last word: Forces for the millennium. Bulletin of the Atomic Scientists, 54(5):72, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:LWMa

[Ark98d] William M. Arkin. The last word: Me and World War III. Bulletin of the Atomic Scientists, 54(3):72, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:LWMb

[Ark98e] William M. Arkin. The last word: Misplaced priorities. Bulletin of the Atomic Scientists, 54(4):72, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1998:LWN

[Ark98f] William M. Arkin. The last word: No points safe. Bulletin of the Atomic Scientists, 54(1):73, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arkin:1999:EE

[Ark99a] William M. Arkin. End of an era. Bulletin of the Atomic Scientists, 55(3):81, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/3/81.

Arkin:1999:PHB

[Ark99b] William M. Arkin. In praise of heavy bombers. Bulletin of the Atomic Scientists, 55(4):80, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/4/80.

Arkin:1999:NIR

[Ark99c] William M. Arkin. A new idea for reductions. Bulletin of the Atomic Scientists, 55(1):81, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/81.

Arkin:1999:OW

[Ark99d] William M. Arkin. One to watch. Bulletin of the Atomic Scientists, 55(5):72, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/5/72.

Arkin:1999:PDM

[Ark99e] William M. Arkin. Politically depleted munitions. Bulletin of the Atomic Scientists, 55(6):72, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/6/72.

Arkin:2000:IM

[Ark00a] William M. Arkin. Iceland melts. Bulletin of the Atomic Scientists, 56(1):80, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/80.

Arkin:2000:NNN

[Ark00b] William M. Arkin. No nukes, or new nukes? Bulletin of the Atomic Scientists, 56(6):84, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/6/84.

Arkin:2000:NB

[Ark00c] William M. Arkin. The nuclear burden. Bulletin of the Atomic Scientists, 56(2):80, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/2/80.

Arkin:2000:RNW

[Ark00d] William M. Arkin. Russian nuclear war, made in Omaha. Bulletin of the Atomic Scientists, 56(4):72, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/4/72.

Arkin:2000:SHW

[Ark00e] William M. Arkin. Saddam Hussein, winner. Bulletin of the Atomic Scientists, 56(3):72, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/3/72.

Arkin:2000:SFI

[Ark00f] William M. Arkin. The SIOP — forever immoral. Bulletin of the Atomic Scientists, 56(5):72, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/5/72.

Arkin:2000:SBD

[Ark00g] William M. Arkin. Smart bombs, dumb targeting? Bulletin of the Atomic Scientists, 56(3):46–53, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/3/46.

Arkin:2001:NP

[Ark01a] William M. Arkin. Nuclear posturing. Bulletin of the Atomic Scientists, 57(3):80, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/3/80.

Arkin:2001:PG

[Ark01b] William M. Arkin. The Praetorian Guards. Bulletin of the Atomic Scientists, 57(2):80, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/80.

Arkin:2001:SL

[Ark01c] William M. Arkin. The silver lining. Bulletin of the Atomic Scientists, 57(1):76, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/1/76.

Arkin:2001:STN

[Ark01d] William M. Arkin. Some things never change. Bulletin of the Atomic Scientists, 57(6):80, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/6/80.

Arkin:2001:TC

[Ark01e] William M. Arkin. Targeting China. Bulletin of the Atomic Scientists, 57(4):72, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/4/72.

Arkin:2001:TD

[Ark01f] William M. Arkin. Theater defense. Bulletin of the Atomic Scientists, 57(5):73, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/5/73.

Arkin:2002:MM

[Ark02a] William M. Arkin. March madness. Bulletin of the Atomic Scientists, 58(3):76, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/3/76.

Arkin:2002:NAE

[Ark02b] William M. Arkin. Numbers aren’t everything. Bulletin of the Atomic Scientists, 58(5):76, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/5/76.

Arkin:2002:WEW

[Ark02c] William M. Arkin. Wanted: Experts on war. Bulletin of the Atomic Scientists, 58(1):72, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/72.

Arkin:2002:WFB

[Ark02d] William M. Arkin. Whose finger on the button? Bulletin of the Atomic Scientists, 58(2):73, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/73.

Arkin:2006:CMF

[Ark06] William M. Arkin. The continuing misuses of fear. Bulletin of the Atomic Scientists, 62(5):42–45, September/October 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/5/42.

Armstrong:1953:LEM

[Arm53] E. L. Armstrong. Letter to the Editor: Moral position of the atomic scientist. Bulletin of the Atomic Scientists, 9(9):353, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Armacost:1968:BBL

[Arm68] Michael H. Armacost. Books: Deadly Logic: The Theory of Nuclear Deterrence, by Philip Green. Bulletin of the Atomic Scientists, 24(2):24–27, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arneson:1947:HFW

[Arn47] R. Gordon Arneson. How far have we gone towards international control of atomic energy?: An estimate of the chances for eventual agreement on a program for international control. Bulletin of the Atomic Scientists, 3(3):83–86, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnold:1950:HCB

[Arn50] James R. Arnold. The hydrogen–cobalt bomb. Bulletin of the Atomic Scientists, 6(10):290–292, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnold:1953:TNE

[Arn53] James R. Arnold. Tracing nuclear explosions. Bulletin of the Atomic Scientists, 9(2):44–45, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1954:LET

[Arn54a] John H. Arnett. Letter to the Editor: The threat from Russia. Bulletin of the Atomic Scientists, 10(3):113, March 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnold:1954:ERB

[Arn54b] James R. Arnold. Effects of the recent bomb tests on human beings. Bulletin of the Atomic Scientists, 10(9):347–348, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnold:1955:FHE

[Arn55] James Arnold. Fall-out hazard: an erratum. Bulletin of the Atomic Scientists, 11(2):52, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnold:1973:CM

[Arn73] James R. Arnold. The Chemist’s moon. Bulletin of the Atomic Scientists, 29(9):22–25, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnold:1980:CPD

[Arn80] Henry F. Arnold. Commentary: Power debate requires humane vision. Bulletin of the Atomic Scientists, 36(9):57, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fai80].

Arnett:1990:RIS

[Arn90] Eric H. Arnett. Report: Iraq: Surgery with a Tomahawk. Bulletin of the Atomic Scientists, 46(9):7, 47, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1991:RGW

[Arn91] Eric H. Arnett. Report: Gulf War: Awestruck press does Tomahawk PR. Bulletin of the Atomic Scientists, 47(3):7–8, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [KA91].

Arnett:1992:BTT

[Arn92a] Eric H. Arnett. Bulletin: Truth and Tomahawks. Bulletin of the Atomic Scientists, 48(6):3–4, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1992:DMS

[Arn92b] Eric H. Arnett. DARPA — many shades of black. Bulletin of the Atomic Scientists, 48(7):21, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1992:RHT

[Arn92c] Eric H. Arnett. A return to the “hair-trigger”. Bulletin of the Atomic Scientists, 48(7):16–17, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1992:WH

[Arn92d] Eric H. Arnett. Welcome to hyperwar. Bulletin of the Atomic Scientists, 48(7):14–21, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1996:RCN

[Arn96a] Eric Arnett. Report: CTB: Nuclear club gets clubbier. Bulletin of the Atomic Scientists, 52(3):12–13, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1996:RSA

[Arn96b] Eric Arnett. Report: South Asia: India’s nuclear brownout. Bulletin of the Atomic Scientists, 52(6):15–16, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dom97].

Arnett:1997:GOS

[Arn97a] Eric Arnett. Guest opinion: And still champion …. Bulletin of the Atomic Scientists, 53(4):4, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1997:WT

[Arn97b] Eric Arnett. What threat? Bulletin of the Atomic Scientists, 53 (2):53–54, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LA98].

Arnett:1998:BSS

[Arn98a] Eric Arnett. Big science, small results. Bulletin of the Atomic Scientists, 54(4):46–47, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1998:LLA

[Arn98b] Eric Arnett. Letter: Look to the army. Bulletin of the Atomic Scientists, 54(5):66–67, September/October 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1998:PII

[Arn98c] Eric Arnett. Perspective: Iran is not Iraq. Bulletin of the Atomic Scientists, 54(1):12–14, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Koc98].

Arnold:2022:SBE

[Arn22] Aaron Arnold. Stolen billions from errant mouse clicks: Crypto requires new approaches to attack money-laundering. Bulletin of the Atomic Scientists, 78(4):191–197, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2087374.

Aron:1950:ABE

[Aro50] Raymond Aron. The atomic bomb and Europe. Bulletin of the Atomic Scientists, 6(4):110–114, 125–126, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aron:1951:APE

[Aro51] Raymond Aron. American policy: Europe and Asia. Bulletin of the Atomic Scientists, 7(3):81–88, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aron:1952:AVP

[Aro52] Raymond Aron. American visa policy. Bulletin of the Atomic Scientists, 8(7):234–235, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Aron:1955:RCS

[Aro55] Raymond Aron. Realism and common sense in security policy. Bulletin of the Atomic Scientists, 11(4):110–112, April 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Aron:1956:HCL

[Aro56] Raymond Aron. A half-century of limited war? Bulletin of the Atomic Scientists, 12(4):99–104, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aron:1960:FWD

[Aro60] Raymond Aron. The future of Western deterrent power. I. A view from France. Bulletin of the Atomic Scientists, 16(7):266– 270, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aron:1964:RMS

[Aro64] Raymond Aron. Rationality of modern society. Bulletin of the Atomic Scientists, 20(1):23–24, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arp:1961:LES

[Arp61] Halton C. Arp. Letter to the Editor: A sky to remember. Bulletin of the Atomic Scientists, 17(10):429, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Artsimovitch:1968:MPC

[Art68] L. A. Artsimovitch. The modern physicist and the case for science politics. Bulletin of the Atomic Scientists, 24(9):23, 41–48, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Artsimovich:1970:C

[Art70] L. A. Artsimovich. Control. Bulletin of the Atomic Scientists, 26 (6):47–55, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Artsimovich:1971:ICL

[Art71] L. A. Artsimovich. II. Comments of Lev Artsimovich, Chief, Soviet Thermonuclear Project. Bulletin of the Atomic Scientists, 27(8):50–51, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arthur:1985:LBI

[Art85] A. D. Arthur. Letter: Bulletin intellectual, trendy. Bulletin of the Atomic Scientists, 41(3):64, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arzumanjan:1961:CD

[Arz61] A. A. Arzumanjan. Comprehensive disarmament. Bulletin of the Atomic Scientists, 17(4):127–129, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrams:1962:SBP

[AS62] Herbert K. Abrams and Samuel Suisman. Shelters and Bulletin policy. Bulletin of the Atomic Scientists, 18(3):38–39, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alvarez:1985:URU

[AS85] Robert Alvarez and Debra Sherman. U.S. to resume uranium production for weapons. Bulletin of the Atomic Scientists, 41(4): 28–30, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Arnett:1994:NHB

[AS94] Eric Arnett and Annette Schaper. No hydronuclear ban. Bulletin of the Atomic Scientists, 50(6):22–23, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Abrahamson:2000:PA

[AS00] Dean Abrahamson and Johan Swahn. The political atom. Bulletin of the Atomic Scientists, 56(4):39–44, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/4/39.

Ackerman:2002:WTI

[AS02] Gary Ackerman and Laura Snyder. Would they if they could? Bulletin of the Atomic Scientists, 58(3):40–47, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/3/40.

Atran:2006:CD

[AS06] Scott Atran and Marc Sageman. Connecting the dots. Bulletin of the Atomic Scientists, 62(4):68, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/4/68.

Ahmad:2016:IME

[AS16] Ali Ahmad and Ryan Snyder. Iran and multinational enrichment in the Middle East. Bulletin of the Atomic Scientists, 72(1):52– 57, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1955:EFb

[ASA55] S. K. Allison, E. Segrè, and Herbert L. Anderson. Enrico Fermi. Bulletin of the Atomic Scientists, 11(1):2, 40, January 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Amati:1984:ICP

[ASB^{+}84] D. Amati, J. Steinberger, Nina Byers, V. F. Weisskopf, R. Hagedorn, and Ch. Wetterich. International call of physicists for an immediate freeze on deployment of new nuclear weapons. Bulletin of the Atomic Scientists, 40(1):37, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ashby:1948:CBD

[Ash48] Eric Ashby. Comments on “Freedom Demands Responsibility”. Bulletin of the Atomic Scientists, 4(10):304–305, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a, Rit48, Ler48, Ano48v].

Ashby:1950:SSU

[Ash50] Eric Ashby. Science in the Soviet Union: A British scientist’s reply. Bulletin of the Atomic Scientists, 6(6):179, June 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ashby:1957:BAT

[Ash57] Eric Ashby. British academic tenure. Bulletin of the Atomic Scientists, 13(4):134, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Asher:1960:LEF

[Ash60] Thomas M. Asher. Letter to the Editor: Five-year armistice for biological warfare. Bulletin of the Atomic Scientists, 16(7):256, 305, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ashkenaz:1976:LDS

[Ash76] Eleanor W. Ashkenaz. Letter: Double standard. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Ashworth:1981:FRB

[Ash81] Lynn W. Ashworth. Film reviews: The Russians, a series of three films People of the Cities, People of the Country, People of Influence. Bulletin of the Atomic Scientists, 37(7):49–50, August/ September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Asimov:1957:LEI

[Asi57] Isaac Asimov. Letter to the Editor: Injustice done? Bulletin of the Atomic Scientists, 13(5):185, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Asimov:1962:BBR

[Asi62a] Isaac Asimov. Books: The Race to the Year 2000, by Fritz Baade. Bulletin of the Atomic Scientists, 18(10):26–27, 29, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Asimov:1962:LES

[Asi62b] Isaac Asimov. Letter to the Editor: Scientist, writer. Bulletin of the Atomic Scientists, 18(7):37–38, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor62b].

Asimov:1963:SA

[Asi63] Isaac Asimov. The sword of Achilles. Bulletin of the Atomic Scientists, 19(9):17–18, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aspin:1978:SNS

[Asp78] Les Aspin. SALT or no SALT. Bulletin of the Atomic Scientists, 34(6):34–38, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aspin:1979:BRB

[Asp79] Les Aspin. Book reviews: Soviet Civil Defense, by the Director of Central Intelligence. Bulletin of the Atomic Scientists, 35(2):44– 45, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aspin:1980:JNA

[Asp80] Les Aspin. Judge not by numbers alone. Bulletin of the Atomic Scientists, 36(6):28–33, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aspin:1983:CMB

[Asp83] Les Aspin. Commentary: The MX bargain. Bulletin of the Atomic Scientists, 39(9):52–54, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Tsi83b, You84a].

Aspin:1987:CFE

[Asp87] Les Aspin. Conventional forces in Europe: Unilateral moves for stability. Bulletin of the Atomic Scientists, 43(10):12–15, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AAU:1953:UTF

[Ass53] Association of American Universities. Universities and their faculties. Bulletin of the Atomic Scientists, 9(5):188–190, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AGS:1963:CDG

[Ass63] Association of German Scientists. Civil defense: A German view. Bulletin of the Atomic Scientists, 19(2):46–48, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AGA:1980:CRW

[Ass80] Association for Geneva Appeal. Commentary: Radioactive waste in Europe. Bulletin of the Atomic Scientists, 36(10):59, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Astiz:1967:CFL

[Ast67] Carlos Alberto Astiz. The changing face of Latin American higher education. Bulletin of the Atomic Scientists, 23(2):4–8, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allardice:1962:FPb

[AT62] Corbin Allardice and Edward R. Trapnell. The first pile. Bulletin of the Atomic Scientists, 18(10):19–24, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1988:MWL

[AT88] David Albright and Theodore B. Taylor. Making warheads: A little tritium goes a long way. Bulletin of the Atomic Scientists, 44(1):39–42, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LH88].

Adams:1992:STS

[AT92] Gordon Adams and Paul Taibl. Share technology for “safer” weapons. Bulletin of the Atomic Scientists, 48(4):38–40, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Atkinson:1978:CRR

[Atk78] Richard C. Atkinson. Commentary: Rights and responsibilities in scientific research. Bulletin of the Atomic Scientists, 34(10):10– 14, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LaM79].

Atlee:1947:PBU

[Atl47] C. R. Atlee. Prominent Britons urge Attlee to call Big Three meeting. Bulletin of the Atomic Scientists, 3(8):218, 220, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Atlee:1954:PP

[Atl54] Clement R. Atlee. The political problem. Bulletin of the Atomic Scientists, 10(8):327–328, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Able:1985:LSW

[ATL85] John Able, Lawrence Trevethan, and Steve Lampman. Letter: On Star Wars editorial. Bulletin of the Atomic Scientists, 41 (8):61–62, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro85i].

AEC:1950:CWA

[Ato50] Atomic Energy Commission. The City of Washington and an atomic attack. Bulletin of the Atomic Scientists, 6(1):29–30, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AEC:1954:ASO

[Ato54a] Atomic Energy Commission. AEC statement [on Oppenheimer case]. Bulletin of the Atomic Scientists, 10(5):187, May 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AEC:1954:RER

[Ato54b] Atomic Energy Commission. Radiation exposures in recent weapons tests. Bulletin of the Atomic Scientists, 10(9):352, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Atran:2006:PBA

[Atr06] Scott Atran. Pakistan: Balancing act. Bulletin of the Atomic Scientists, 62(6):23–24, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/6/23.

Attlee:1954:MBR

[Att54] Clement R. Attlee. [More on the British reaction]. Bulletin of the Atomic Scientists, 10(5):167, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

AuCoin:1984:PF

[AuC84] Les AuCoin. Perspective: Freeze. Bulletin of the Atomic Scientists, 40(9):7–9, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cas85b].

Audrieth:1965:LES

[Aud65] L. F. Audrieth. Letter to the Editor: Scientists on tap or on top? We are missing the real issues. Bulletin of the Atomic Scientists, 21(7):24–25, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Auerbach:1978:BRB

[Aue78] Lewis Auerbach. Book reviews: Autonomous Technology: Technics Out of Control as a Theme in Political Thought, by Langdon Winner. Bulletin of the Atomic Scientists, 34(9):48–49, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Auer:2001:VI

[Aue01] Catherine Auer. A view from inside. Bulletin of the Atomic Scientists, 57(2):69–70, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/69.

Auer:2002:NSU

[Aue02] Catherine Auer. Nuclear safety: Uh-oh in Ohio. Bulletin of the Atomic Scientists, 58(4):14–17, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/4/14.

Auer:2003:AAP

[Aue03a] Catherine Auer. As American as atomic pie. Bulletin of the Atomic Scientists, 59(2):6–8, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

www.highbeam.com/doc/1G1-99147968.html.

Auer:2003:AW

[Aue03b] Catherine Auer. Atoms for what? Bulletin of the Atomic Scientists, 59(6):42–43, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/6/42.

Auer:2003:IAJ

[Aue03c] Catherine Auer. Is it all just a game? Bulletin of the Atomic Scientists, 59(4):36–37, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/4/36.

Auer:2003:KNL

[Aue03d] Catherine Auer. Killer ‘non-lethals’. Bulletin of the Atomic Scientists, 59(1):42, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Auer:2004:IM

[Aue04] Catherine Auer. In it for the money. Bulletin of the Atomic Scientists, 60(2):42–43, March/April 2004. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/2/42.

Auer:2005:NME

[Aue05a] Catherine Auer. No more excuses. Bulletin of the Atomic Scientists, 61(6):69, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/6/69.

Auer:2005:RW

[Aue05b] Catherine Auer. Received wisdom. Bulletin of the Atomic Scientists, 61(4):70, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/4/70.

Auer:2006:BHF

[Aue06a] Catherine Auer. Better homes and fallout shelters. Bulletin of the Atomic Scientists, 62(3):58–59, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/3/58.

Auer:2006:FV

[Aue06b] Catherine Auer. Finding a voice. Bulletin of the Atomic Scientists, 62(1):65–66, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/1/65.2.

Auer:2006:RSK

[Aue06c] Catherine Auer. A real skeleton key. Bulletin of the Atomic Scientists, 62(1):36–37, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/1/36.

Augur:1948:DCF

[Aug48a] Tracy B. Augur. The dispersal of cities — a feasible program. Bulletin of the Atomic Scientists, 4(10):312–315, October 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Augur:1948:DCD

[Aug48b] Tracy B. Augur. The dispersal of cities as a defense measure. Bulletin of the Atomic Scientists, 4(5):131–134, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Augur:1948:NSF

[Aug48c] Tracy B. Augur. National security factors in industrial location. Bulletin of the Atomic Scientists, 4(10):316–317, 320, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Augur:1950:DGB

[Aug50] Tracy B. Augur. Dispersal is good business. Bulletin of the Atomic Scientists, 6(8–9):244–245, August/September 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Auger:1955:WWH

[Aug55] Pierre Auger. Who? Why? How? Bulletin of the Atomic Scientists, 11(3):74–76, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Nie55, SD55].

Auger:1956:LEF

[Aug56a] Pier Auger. Letter to the Editor: The founding of CERN. Bulletin of the Atomic Scientists, 12(6):232, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Auger:1956:PIS

[Aug56b] Pierre Auger. Part III: Science and social responsibility: Science as a force for unity among men. Bulletin of the Atomic Scientists, 12(6):208–210, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Auger:1965:LS

[Aug65] Pierre Auger. Limits to science. Bulletin of the Atomic Scientists, 21(9):21–22, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurand:1946:ARP

[Aur46] H. S. Aurand. The Army’s research program. Bulletin of the Atomic Scientists, 2(9–10):10, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLF

[Aur71a] Aurum. Letter from London: A funereal occasion. Bulletin of the Atomic Scientists, 27(4):29–30, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLC

[Aur71b] Aurum. Letter from London: Changing the guard. Bulletin of the Atomic Scientists, 27(2):42–43, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLD

[Aur71c] Aurum. Letter from London: Development and the Kalinga Prize. Bulletin of the Atomic Scientists, 27(3):24, 26–27, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLM

[Aur71d] Aurum. Letter from London: Myrdal, Benjamin Franklin and scienomics. Bulletin of the Atomic Scientists, 27(9):31–32, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLN

[Aur71e] Aurum. Letter from London: Notes on science policy in Britain. Bulletin of the Atomic Scientists, 27(10):21–22, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLS

[Aur71f] Aurum. Letter from London: Science policy and the question of relevancy. Bulletin of the Atomic Scientists, 27(6):25–26, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1971:LLT

[Aur71g] Aurum. Letter from London: Technology and democratic performance. Bulletin of the Atomic Scientists, 27(5):41–42, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1972:LLC

[Aur72a] Aurum. Letter from London: A challenge for the prophets of calamity. Bulletin of the Atomic Scientists, 28(5):28, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1972:LLBa

[Aur72b] Aurum. Letter from London: Britain chooses a reactor. Bulletin of the Atomic Scientists, 28(2):45–46, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1972:LLBc

[Aur72c] Aurum. Letter from London: Britain’s research establishments. Bulletin of the Atomic Scientists, 28(4):43–44, April 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1972:LLBb

[Aur72d] Aurum. Letter from London: British architects and the environment. Bulletin of the Atomic Scientists, 28(3):42–43, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1972:LLP

[Aur72e] Aurum. Letter from London: Population policy in Britain. Bulletin of the Atomic Scientists, 28(1):26–27, January 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Aurum:1972:LLR

[Aur72f] Aurum. Letter from London: Report from CERN. Bulletin of the Atomic Scientists, 28(9):44, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Austin:1947:OSO

[Aus47] Warren R. Austin. Our states and the opportunity for peace and security. Bulletin of the Atomic Scientists, 3(2):59–61, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Austern:1960:LES

[Aus60] Norman Austern. Letter to the Editor: Science’s role granted. Bulletin of the Atomic Scientists, 16(1):33, January 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Doo59].

Austin:1981:CEA

[Aus81] Joseph Austin. Commentary: Educating for annihilation? Bulletin of the Atomic Scientists, 37(4):65, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shu81].

Austin:1999:HD

[Aus99] Kathi Austin. Hearts of darkness. Bulletin of the Atomic Scientists, 55(1):34–37, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/34.

Aviel:1983:CCV

[Avi83] David Aviel. Commentary: Congressional views on nuclear energy. Bulletin of the Atomic Scientists, 39(2):48, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Alexander:1984:E

[AW84] Susan Alexander and Tony Wagner. Education. Bulletin of the Atomic Scientists, 40(10):20s–21s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adams:1985:MSB

[AW85] Gordon Adams and Laura Weiss. Military spending boosts the deficit. Bulletin of the Atomic Scientists, 41(4):26–27, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1950:NB

[Axe50] Peter Axel. News in brief. Bulletin of the Atomic Scientists, 6 (12):380–382, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:MD

[Axe51a] Peter Axel. Manpower and deferment. Bulletin of the Atomic Scientists, 7(5):131–132, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NNa

[Axe51b] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 7(6):189–191, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NNb

[Axe51c] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 7 (7–8):239–240, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NNc

[Axe51d] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 7 (10):315–317, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NNd

[Axe51e] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 7(11):350–352, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NNe

[Axe51f] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 7 (12):380–382, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NNM

[Axe51g] Peter Axel. News and notes: Military and industrial use of fission products. Bulletin of the Atomic Scientists, 7(5):153–157, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NBb

[Axe51h] Peter Axel. News in brief. Bulletin of the Atomic Scientists, 7(1): 30–32, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NBc

[Axe51i] Peter Axel. News in brief. Bulletin of the Atomic Scientists, 7 (2):62–64, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NBd

[Axe51j] Peter Axel. News in brief. Bulletin of the Atomic Scientists, 7(3): 95–96, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1951:NBA

[Axe51k] Peter Axel. News in brief: The armament debate: Land armies or atomic air force? Bulletin of the Atomic Scientists, 7(4):123–124, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1952:NNa

[Axe52a] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8(1):31–32, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1952:NNb

[Axe52b] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8(2):60–61, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1952:NNc

[Axe52c] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8(3):94–95, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1952:NNd

[Axe52d] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8(4):127–128, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1952:NNe

[Axe52e] Peter Axel. News and notes. Bulletin of the Atomic Scientists, 8(5):165–168, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axel:1960:SEV

[Axe60] Peter Axel. Scientific exchange visits. Bulletin of the Atomic Scientists, 16(6):212–215, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Axtmann:1973:PSC

[Axt73] Robert C. Axtmann. Padre Sol and chemical fuel production. Bulletin of the Atomic Scientists, 29(8):42–44, October 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ayensu:1985:ATF

[Aye85] Edward S. Ayensu. Africa: Toward food self-sufficiency. Bulletin of the Atomic Scientists, 41(8):50–52, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1989:BVH

[AZ89a] David Albright and Tom Zamora. Bulletin: Valindaba humming. Bulletin of the Atomic Scientists, 45(8):5–6, October 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1989:IPN

[AZ89b] David Albright and Tom Zamora. India, Pakistan’s nuclear weapons: all the pieces in place. Bulletin of the Atomic Scientists, 45(5):20–26, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1990:WP

[AZ90] David Albright and Tom Zamora. Warheads in the pipeline. Bulletin of the Atomic Scientists, 46(5):18, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1991:SAF

[AZ91] David Albright and Tom Zamora. South Africa flirts with the NPT. Bulletin of the Atomic Scientists, 47(1):27–31, January/ February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Azhgikhina:1993:HCM

[Azh93] Nadezhda Azhgikhina. High culture meets trash TV. Bulletin of the Atomic Scientists, 49(1):42–46, January/February 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Azhgikhina:1995:MWW

[Azh95a] Nadezhda Azhgikhina. More women’s work. Bulletin of the Atomic Scientists, 51(4):50, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Azhgikhina:1995:MB

[Azh95b] Nadezhda Azhgikhina. A movement is born. Bulletin of the Atomic Scientists, 51(4):47–53, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translation from Russian by Viktoria Tripolskaya-Mitlyng.

Albright:1990:PT

[AZL90a] David Albright, Tom Zamora, and David Lewis. Plutonium by the ton. Bulletin of the Atomic Scientists, 46(5):15, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Albright:1990:TRF

[AZL90b] David Albright, Tom Zamora, and David Lewis. Turn off Rocky Flats. Bulletin of the Atomic Scientists, 46(5):12–16, 18–19, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bogart:1989:LBU

[BA89] Charles H. Bogart and Len Ackland. Letter: Bulletin unbalanced. Bulletin of the Atomic Scientists, 45(2):48, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bielefeld:2007:KJ

[BA07] Tom Bielefeld and Hassan Abbas. The Kahn job. Bulletin of the Atomic Scientists, 63(4):72–73, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/72.

Bethe:1975:SSN

[BAA^{+}75] Hans A. Bethe, Luis Alvarez, Peter Auer, William O. Baker, John Bardeen, Robert F. Bacher, Felix Bloch, Norris E. Bradbury, Harold Brown, Richard H. Chamberlain, Cyril L. Comar, Arthur Kantrowitz, Ralph E. Lapp, Joshua Lederberg, Willard F. Libby, Franklin A. Long, Edwin M. McMillan, Kenneth S. Pitzer, Edward M. Purcell, I. I. Rabi, Norman Rasmussen, Roger Revelle, Glenn T. Seaborg, Frederick Seitz, Edward Teller, James A. Van Allen, Warren Weaver, Alvin Weinberg, Victor F. Weisskopf, Edward Wenk, Jr., Eugene Wigner, and Richard Wilson. 32 scientists speak out: “No alternative to nuclear power”. Bulletin of the Atomic Scientists, 31(3):4–5, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Wei75] about the incorrect inclusion of Victor Weisskopf in the authors. See comments [Gir75, Smi75d, Sau75, Gul75].

Benedict:1946:TCA

[BAB^{+}46] M. Benedict, L. W. Alvarez, L. A. Bliss, S. G. English, A. B. Kinzell, P. Morrison, F. H. English, C. Starr, and W. J. Williams.

Technological control of atomic energy activities: Report based on studies by American scientists — transmitted to the UN Atomic Energy Commission by the US Delegation. Bulletin of the Atomic Scientists, 2(11–12):18–29, December 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Babcock:1993:RCP

[Bab93] Glenys A. Babcock. Report: Chernobyl: Perceptions of peril. Bulletin of the Atomic Scientists, 49(5):11–12, June 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacher:1948:PFD

[Bac48] Robert F. Bacher. The physicist and the future development of atomic energy. Bulletin of the Atomic Scientists, 4(4):99–102, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacher:1949:DNR

[Bac49] Robert F. Bacher. The development of nuclear reactors. Bulletin of the Atomic Scientists, 5(3):80–82, 94, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacher:1950:HB

[Bac50] Robert F. Bacher. The hydrogen bomb. Bulletin of the Atomic Scientists, 6(5):133–138, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacq:1958:RMB

[Bac58] Z. M. Bacq. Radiation and man: Broad studies: The U.N. Radiation Committee. Bulletin of the Atomic Scientists, 14(1):55–56, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bachrach:1959:ET

[Bac59] Arthur J. Bachrach. The ethics of tachistoscopy. Bulletin of the Atomic Scientists, 15(5):212–215, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacher:1977:NEO

[Bac77a] Robert F. Bacher. Nuclear energy and our future. Bulletin of the Atomic Scientists, 33(3):63–65, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bachmaier:1977:LIC

[Bac77b] James T. Bachmaier. Letter: An informal consensus. Bulletin of the Atomic Scientists, 33(2):6, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacher:1980:TFS

[Bac80] Robert F. Bacher. A time to find solutions. Bulletin of the Atomic Scientists, 36(2):46–47, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacher:1981:ERN

[Bac81] Robert F. Bacher. Editorial: On the reduction of nuclear weapons. Bulletin of the Atomic Scientists, 37(9):1, 4–5, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bacon:2003:IHC

[Bac03] Kenneth H. Bacon. Iraq: The humanitarian challenge. Bulletin of the Atomic Scientists, 59(1):26–27, January/February 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/1/26.

Backus:2012:AWC

[Bac12] George Backus. Arctic 2030: What are the consequences of climate change?: The US response. Bulletin of the Atomic Scientists, 68(4):9–16, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/68/4/9.

Badeau:1966:DDM

[Bad66] John S. Badeau. Development and diplomacy in the Middle East. Bulletin of the Atomic Scientists, 22(5):5–10, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Badash:1978:WPC

[Bad78] Lawrence Badash. The worst possible case. Bulletin of the Atomic Scientists, 34(7):61–62, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [CB79].

Baer:1960:LET

[Bae60] R. M. Baer. Letter to the Editor: Tests engender mistrust. Bulletin of the Atomic Scientists, 16(3):80, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bainbridge:1975:AOTb

[Bai75a] K. T. Bainbridge. ‘All in our time’: A foul and awesome display. Bulletin of the Atomic Scientists, 31(5):40–46, May 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bainbridge:1975:AOTa

[Bai75b] Kenneth T. Bainbridge. ‘All in our time’: Prelude to Trinity. Bulletin of the Atomic Scientists, 31(4):42–46, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bailey:1980:WWW

[Bai80] Kathleen Bailey. When and why weapons. Bulletin of the Atomic Scientists, 36(4):42–45, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bailey:1992:RSA

[Bai92] Greg Bailey. Report: Strategic Air Command: Farewell to SAC. Bulletin of the Atomic Scientists, 48(5):10–11, June 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bailey:1995:NWW

[Bai95] Kathleen Bailey. Nonproliferation: Why we have to keep the bomb. Bulletin of the Atomic Scientists, 51(1):30–37, January/ February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [RGK95].

Bailey:1996:W

[Bai96] Kathleen Bailey. So what? Bulletin of the Atomic Scientists, 52(5):46–47, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bainbridge:2020:AOT

[Bai20] Kenneth T. Bainbridge. 1975: All in our time: A foul and awesome display. Bulletin of the Atomic Scientists, 76(6):374–380, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847493.

Bakker:1960:CII

[Bak60] C. J. Bakker. CERN as an institute for international cooperation. Bulletin of the Atomic Scientists, 16(2):54–57, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baker:1975:TMP

[Bak75] Jeffrey J. W. Baker. Three modes of protest action: The search for ways of mobilizing scientists in the public affairs area. Bulletin of the Atomic Scientists, 31(2):8–15, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baker:1978:MAN

[Bak78] Philip Noel Baker. A message for all nations. Bulletin of the Atomic Scientists, 34(5):9, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baker:1986:LTD

[Bak86] Dorothy G. Baker. Letter: Truman’s decision. Bulletin of the Atomic Scientists, 42(1):56–57, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mes85b].

Baker:2006:SE

[Bak06] Robert J. Baker. Surprise ending. Bulletin of the Atomic Scientists, 62(1):59–61, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/1/59.

Baldwin:1950:HBP

[Bal50] Hanson W. Baldwin. Hanson Baldwin on preventive war. Bulletin of the Atomic Scientists, 6(10):318, October 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the author’s article in the New York Times, 1 September 1950.

Baldwin:1954:NPN

[Bal54] Hanson W. Baldwin. Nagasaki plus nine years. Bulletin of the Atomic Scientists, 10(8):318, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baldwin:1956:NFW

[Bal56] Hanson W. Baldwin. The new face of war. Bulletin of the Atomic Scientists, 12(5):153–158, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baldwin:1960:RCT

[Bal60a] Hanson Baldwin. A reason for continuing tests. Bulletin of the Atomic Scientists, 16(5):172, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Two-paragraph excerpt from the author’s article in the 3 March 1960 issue of the New York Times.

Ball:1960:LED

[Bal60b] Albert P. Ball. Letter to the Editor: “Delusion” author deluded? Bulletin of the Atomic Scientists, 16(1):0, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Balio:1967:PCH

[Bal67] Tino T. Balio. The public confrontation of Hermann J. Muller. Bulletin of the Atomic Scientists, 23(9):8–12, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Balint:1976:LMS

[Bal76] David J. Balint. Letter: Move over, Spiro. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Baltimore:1978:CRD

[Bal78] David Baltimore. Commentary: Recombinant DNA research. Bulletin of the Atomic Scientists, 34(10):56–57, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gre78].

Balch:1979:PNP

[Bal79a] Emily Greene Balch. Perspective: Nobel Peace Laureates: 1946. Bulletin of the Atomic Scientists, 35(4):1, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 7 April 1948, Oslo, Norway.

Ball:1979:BRBb

[Bal79b] Nicole Ball. Book review: The Twenty-Ninth Day: Accommodating Human Needs and Numbers to the Earth’s Resources, by Lester R. Brown. Bulletin of the Atomic Scientists, 35(9):51– 52, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ball:1979:BRBa

[Bal79c] Nicole Ball. Book reviews: The Atmosphere: Endangered and Endangering, edited by William W. Kellogg and Margaret Mead; Earth’s Aura, by Louise B. Young; A Change in the Weather, Fitzhugh Green; Regimes for the Ocean, Outer Space and Weather, by Seyom Brown, Nina W. Cornell, Larry L. Fabian, and Edith Brown Weiss. Bulletin of the Atomic Scientists, 35(4):34–36, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Balzer:1982:SUS

[Bal82] Harley D. Balzer. The Soviet Union: Scientific and technical education. Bulletin of the Atomic Scientists, 38(7):24–29, August/ September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ball:1984:FPT

[Bal84a] George Ball. Foreign policy: a tragedy of errors. Bulletin of the Atomic Scientists, 40(6):4–7, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ric84].

Ball:1984:PSA

[Bal84b] George W. Ball. Perspective: Slaves to animosity. Bulletin of the Atomic Scientists, 40(7):5–7, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mac84a, Mot85, Sea85b].

Ball:1984:PPN

[Bal84c] George W. Ball. Perspective: The President’s nuclear responsibility. Bulletin of the Atomic Scientists, 40(9):5–7, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ball:1985:EUF

[Bal85] George W. Ball. Erosion of U.S. foreign relations. Bulletin of the Atomic Scientists, 41(7):110–113, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Balazs:1986:ASW

[Bal86a] Louis A. P. Balazs. Anti-Star Wars drive. Bulletin of the Atomic Scientists, 42(3):55–56, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [KW86].

Balzas:1986:LWL

[Bal86b] Lois A. P. Balzas. Letter: Weapons labs and taxpayers. Bulletin of the Atomic Scientists, 42(10):51–52, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Van86].

Balazs:1987:LON

[Bal87] Louis A. P. Balázs. Letter: OELs and NOEMs in the sky. Bulletin of the Atomic Scientists, 43(9):52, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gar87].

Balzer:1988:BRB

[Bal88] Harley Balzer. Book review: The Communist Party and Soviet Science, by Stephen Fortescue. Bulletin of the Atomic Scientists, 44(2):48, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Balkwill:1989:LOJ

[Bal89a] Jack Balkwill. Letter: Outstanding job. Bulletin of the Atomic Scientists, 45(10):50, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ball:1989:ACC

[Bal89b] Desmond Ball. Arms control criteria: A sad record. Bulletin of the Atomic Scientists, 45(4):41–42, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Balakrishnan:1993:LTT

[Bal93] M. R. Balakrishnan. Letter: Think thorium. Bulletin of the Atomic Scientists, 49(9):55, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Buk93].

Bamford:1985:BRB

[Bam85] James Bamford. Book review: The Missing Dimension: Governments and Intelligence Communities in the Twentieth Century, edited by Christopher Andrew and David Dilks. Bulletin of the Atomic Scientists, 41(9):51–52, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bamford:2000:SS

[Bam00] James Bamford. Spies in the sky. Bulletin of the Atomic Scientists, 56(1):65–66, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/65.

Bansak:2011:MNR

[Ban11] Kirk C. Bansak. Managing networks of risk: a tailored approach to Iran’s biological warfare threat potential. Bulletin of the Atomic Scientists, 67(5):66–76, September/October 2011. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/67/5/66.

Banafa:2022:MHC

[Ban22] Ahmed Banafa. Microchips in humans: Consumer-friendly app, or new frontier in surveillance? Bulletin of the Atomic Scientists, 78(5):256–260, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2109330.

Baruch:1946:API

[Bar46a] Bernard M. Baruch. The American proposal for international control. Bulletin of the Atomic Scientists, 2(1–2):3–5, 10, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baruch:1946:MP

[Bar46b] Bernard M. Baruch. Memorandum to the President. Bulletin of the Atomic Scientists, 2(7–8):4–5, 31, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barron:1948:NMR

[Bar48] E. S. Guzman Barron. The Nation’s medical research. Bulletin of the Atomic Scientists, 4(2):59–60, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barron:1957:WDS

[Bar57] Arthur S. Barron. Why do scientists read science fiction? Bulletin of the Atomic Scientists, 13(2):62–65, 70, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barbaclough:1958:BBR

[Bar58a] Geoffrey Barbaclough. Books: German Rearmament and Atomic War, by Hans Speier. Bulletin of the Atomic Scientists, 14(4):153, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnes:1958:QQ

[Bar58b] Sir George Barnes. Quality or quantity? Bulletin of the Atomic Scientists, 14(9):380–384, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baranson:1960:NPS

[Bar60] Jack Baranson. National programs for science and technology in the underdeveloped areas. Bulletin of the Atomic Scientists, 16(5):151–154, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baranson:1961:NTD

[Bar61a] Jack Baranson. New technologies for developing areas. Bulletin of the Atomic Scientists, 17(7):275–277, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barden:1961:LEC

[Bar61b] John Barden. Letter to the Editor: Comments on biological warfare. Bulletin of the Atomic Scientists, 17(2):81, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barden:1963:FUA

[Bar63] John Barden. Forum: The unhappy arbiters of values, standards, and tastes. Bulletin of the Atomic Scientists, 19(10):34–36, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Man64] and response [Bar64a].

Barden:1964:BYN

[Bar64a] John A. Barden. Barden yields to no Mann. Bulletin of the Atomic Scientists, 20(8):31–32, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar63, Man64].

Barnes:1964:LEA

[Bar64b] Wyatt E. Barnes. Letter to the Editor: The atom. Bulletin of the Atomic Scientists, 20(2):25–26, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lil63].

Barden:1965:LEV

[Bar65] John Barden. Letter to the Editor: On Vietnam. Bulletin of the Atomic Scientists, 21(7):29–31, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barden:1966:BBC

[Bar66a] John Barden. Books: Science and Culture, A Study of Cohesive and Disjunctive Force, edited by Gerald Holton. Bulletin of the Atomic Scientists, 22(3):28–30, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnes:1966:LAF

[Bar66b] Peter Barnes. Latin America: The first nuclear free zone? Bulletin of the Atomic Scientists, 22(10):37–40, December 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barber:1967:BBS

[Bar67a] Willard F. Barber. Books: The Sandino Affair, by Neill Macauley. Bulletin of the Atomic Scientists, 23(6):56–57, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnes:1967:BBD

[Bar67b] Peter Barnes. Books: The Denuclearization of Latin America, by Alfonso García Robles, translated by Marjorie Urquidi. Bulletin of the Atomic Scientists, 23(9):32–33, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barton:1967:CRC

[Bar67c] John Barton. Comments: A reconsideration of the criteria for deterrence. Bulletin of the Atomic Scientists, 23(10):41–44, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnes:1969:FAA

[Bar69] N. Spencer Barnes. The Frankenstein argument against peacekeeping. Bulletin of the Atomic Scientists, 25(5):16–18, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barber:1973:BBP

[Bar73a] Willard Barber. Books: Nuclear Politics: The British Experience with an Independent Strategic Force, 1939–1970, by Andrew J. Pierre. Bulletin of the Atomic Scientists, 29(6):44–46, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Pie73].

Barus:1973:CEL

[Bar73b] Carl Barus. Communications: Energy limits. Bulletin of the Atomic Scientists, 29(2):4, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WH72].

Barnea:1974:LWE

[Bar74] Joseph Barnea. Letter: On world energy strategies. Bulletin of the Atomic Scientists, 30(9):2, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lov74b, Lov74a].

Barack:1976:BL

[Bar76a] Rabbi Nathan A. Barack. ‘A brief for licentiousness’. Bulletin of the Atomic Scientists, 32(3):61–62, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pre75] and comment [Pre76].

Barack:1976:UA

[Bar76b] Rabbi Nathan A. Barack. Unwitting accomplices. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Barnaby:1976:EW

[Bar76c] Frank Barnaby. Environmental warfare. Bulletin of the Atomic Scientists, 32(5):36–43, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1976:WAD

[Bar76d] Frank Barnaby. World armament & disarmament. Bulletin of the Atomic Scientists, 32(6):25–32, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barkenbus:1977:NEB

[Bar77a] Jack N. Barkenbus. Nuclear energy and the ballot. Bulletin of the Atomic Scientists, 33(4):4–5, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1977:CBC

[Bar77b] Frank Barnaby. The continuing body count at Hiroshima and Nagasaki: the death and injury toll grows larger, but greater effort will be required to determine the effects more precisely. Bulletin of the Atomic Scientists, 33(10):48–51, 53, December 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste78].

Barnaby:1977:MPN

[Bar77c] Frank Barnaby. The mounting prospects of nuclear war: a report of the Stockholm International Peace Research Institute. Bulletin of the Atomic Scientists, 33(6):10–20, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hel77b].

Barnett:1977:A

[Bar77d] Tim P. Barnett. An allegory. Bulletin of the Atomic Scientists, 33 (2):39, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baranson:1978:CPW

[Bar78a] Jack Baranson. The cornucopian politics of world development. Bulletin of the Atomic Scientists, 34(9):41–44, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1978:WAR

[Bar78b] Frank Barnaby. World arsenals in 1977: a report of the Stockholm International Peace Research Institute. Bulletin of the Atomic Scientists, 34(5):10–13, 45–50, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnhill:1978:LWT

[Bar78c] Bruce M. Barnhill. Letter: War toys. Bulletin of the Atomic Scientists, 34(3):2, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Koe77].

Barnaby:1979:ARS

[Bar79a] Frank Barnaby. Annual report from the Stockholm International Peace Research Institute: World arsenals in 1978. Bulletin of the Atomic Scientists, 35(7):18–26, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1979:BRB

[Bar79b] Frank Barnaby. Book review: The Politics of Peril: Economics, Society, and War, by Carl Friedrich von Weiszäcker. Bulletin of the Atomic Scientists, 35(10):38, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1979:FET

[Bar79c] Wendy Barnaby. First the election and then the referendum. Bulletin of the Atomic Scientists, 35(9):43–47, November 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1979:PRB

[Bar79d] Wendy Barnaby. Play review: Stormen [The Storm]. Bulletin of the Atomic Scientists, 35(5):38–39, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barkenbus:1980:WT

[Bar80a] Jack N. Barkenbus. Whither the treat? Bulletin of the Atomic Scientists, 36(4):37–39, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:ARS

[Bar80b] Frank Barnaby. Annual report of the Stockholm International Peace Research Institute: World arsenals in 1980. Bulletin of the Atomic Scientists, 36(7):9–14, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CGN

[Bar80c] Frank Barnaby. Commentary from Geneva: The NPT review conference — much talk, few results. Bulletin of the Atomic Scientists, 36(9):7–8, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CJF

[Bar80d] Frank Barnaby. Commentary from Jerusalem: A frangible peace. Bulletin of the Atomic Scientists, 36(8):6–7, October 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSC

[Bar80e] Frank Barnaby. Commentary from Stockholm: CBW — an unresolved horror. Bulletin of the Atomic Scientists, 36(6):8–10, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSM

[Bar80f] Frank Barnaby. Commentary from Stockholm: Maneuvers in the Indian Ocean. Bulletin of the Atomic Scientists, 36(5):8–11, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSN

[Bar80g] Frank Barnaby. Commentary from Stockholm: ‘Nuclear explosions are commonplace’. Bulletin of the Atomic Scientists, 36(4): 9–11, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSRa

[Bar80h] Frank Barnaby. Commentary from Stockholm: Reconciling war and humanity. Bulletin of the Atomic Scientists, 36(2):5–6, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSW

[Bar80i] Frank Barnaby. Commentary from Stockholm: War-fighting weapons for Europe. Bulletin of the Atomic Scientists, 36(3): 8–10, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSI

[Bar80j] Frank Barnaby. Commentary: Scandinavian initiatives. Bulletin of the Atomic Scientists, 36(10):5–6, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:SNR

[Bar80k] Frank Barnaby. On the second NPT review conference. Bulletin of the Atomic Scientists, 36(7):20–23, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:CSRb

[Bar80l] Wendy Barnaby. Commentary: The Swedish referendum: ‘Do away with it but not yet’. Bulletin of the Atomic Scientists, 36 (6):58–59, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1980:FRB

[Bar80m] Wendy Barnaby. Film review: Nuclear Nightmares: The Wars That Must Never Happen, produced by Peter Batty. Bulletin of the Atomic Scientists, 36(4):52–53, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1981:ARS

[Bar81a] Frank Barnaby. Annual report of the Stockholm International Peace Research Institute: World arsenals in 1981. Bulletin of the Atomic Scientists, 37(7):16–20, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sno81].

Barnaby:1981:CMN

[Bar81b] Frank Barnaby. Commentary from Moscow: A nuclear engineer’s paradise. Bulletin of the Atomic Scientists, 37(2):55–56, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Coc81a].

Barnaby:1981:CSA

[Bar81c] Frank Barnaby. Commentary from Stockholm: Arms industry — a seller’s market. Bulletin of the Atomic Scientists, 37(5):10–12, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1981:CSC

[Bar81d] Frank Barnaby. Commentary from Stockholm: CO_{2}: Proceed with caution. Bulletin of the Atomic Scientists, 37(4):10–11, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ids81].

Barnaby:1981:CSE

[Bar81e] Frank Barnaby. Commentary from Stockholm: Europeans want to be counted. Bulletin of the Atomic Scientists, 37(3):6–7, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1981:CSM

[Bar81f] Frank Barnaby. Commentary from Stockholm: Militaryscientists. Bulletin of the Atomic Scientists, 37(6):11–12, June/ July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DeW81].

Barnaby:1981:CSV

[Bar81g] Frank Barnaby. Commentary from Stockholm: Victimless secrets. Bulletin of the Atomic Scientists, 37(1):56–58, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1981:CSN

[Bar81h] Wendy Barnaby. Commentary from Stockholm: The nuclear industry and its public. Bulletin of the Atomic Scientists, 37(5): 56–57, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1981:IFS

[Bar81i] Wendy Barnaby. International Foundation for Science: Aid for young scientists. Bulletin of the Atomic Scientists, 37(9):26–27, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1982:CUK

[Bar82a] Frank Barnaby. Commentary from the United Kingdom: United Nations Centre for Disarmament. Bulletin of the Atomic Scientists, 38(1):6–7, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1982:EA

[Bar82b] Frank Barnaby. Europe aroused. Bulletin of the Atomic Scientists, 38(2):8–9, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnaby:1982:FF

[Bar82c] Frank Barnaby. The Falklands fallout. Bulletin of the Atomic Scientists, 38(7):34–35, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baranovsky:1983:MSP

[Bar83a] Vladimir Baranovsky. From Moscow: The search for peace in Europe. Bulletin of the Atomic Scientists, 39(2):10–12, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barash:1983:CWA

[Bar83b] David P. Barash. Commentary: Who’s afraid of the big bad bomb? Bulletin of the Atomic Scientists, 39(1):51–52, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnet:1984:AAA

[Bar84] Richard J. Barnet. Agenda 1984: The Atlantic Alliance. Bulletin of the Atomic Scientists, 40(1):8–10, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barney:1985:LGF

[Bar85a] Dick Barney. Letter: Global future. Bulletin of the Atomic Scientists, 41(2):51, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim84d].

Barton:1985:SBP

[Bar85b] David Barton. Superpowers block progress at Stockholm. Bulletin of the Atomic Scientists, 41(3):36–38, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bor85].

Barton:1985:ACY

[Bar85c] John H. Barton. Arms control and the year 2000. Bulletin of the Atomic Scientists, 41(7):164–167, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnett:1987:ARS

[Bar87] Roger W. Barnett. Arms race at sea: U.S. maritime strategy: sound and safe. Bulletin of the Atomic Scientists, 43(7):30–33, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnet:1988:BRB

[Bar88] Richard J. Barnet. Book review: A Grand Strategy for the West, by Helmut Schmidt. Bulletin of the Atomic Scientists, 44(3): 50–51, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnet:1990:UIL

[Bar90a] Richard J. Barnet. U.S. intervention: Low-intensity thinking. Bulletin of the Atomic Scientists, 46(4):34–37, May 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnouw:1990:FRP

[Bar90b] Erik Barnouw. Film review: PBS takes control: Losing Control?, produced by Gary Krane and Gary Houser. Bulletin of the Atomic Scientists, 46(5):40, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnouw:1990:FRR

[Bar90c] Erik Barnouw. Film review: Radioactive turtles: Building Bombs, produced by Mark Mori and Susan J. Robinson. Bulletin of the Atomic Scientists, 46(4):49–50, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnet:1992:TAN

[Bar92a] Richard J. Barnet. Twin anachronisms: Nuclear weapons and militarism. Bulletin of the Atomic Scientists, 48(4):26–27, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barrillot:1992:FFN

[Bar92b] Bruno Barrillot. French finesse nuclear future. Bulletin of the Atomic Scientists, 48(7):22–26, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barletta:1998:RCM

[Bar98] Michael Barletta. Report: Cyprus: Mediterranean countdown. Bulletin of the Atomic Scientists, 54(6):12–14, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barak:2010:GME

[Bar10a] Eitan Barak. Getting the Middle East holdouts to join the CWC. Bulletin of the Atomic Scientists, 66(1):57–62, January 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/1/57.

Barzegar:2010:IQH

[Bar10b] Kayhan Barzegar. The Iranian quagmire: How to move forward: Position: Tit-for-tat diplomacy. Bulletin of the Atomic Scientists, 66(6):109–114, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/66/6/109.

Basalla:1968:SSS

[Bas68] George Basalla. Science, society, and science education. Bulletin of the Atomic Scientists, 24(6):45–48, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [GS68].

Basile:1980:BRE

[Bas80] Paul S. Basile. Book reviews: Energy soothsayers not saying the whole sooth: Energy: The Next Twenty Years; Nuclear Power: Issues and Choices; Nuclear or Not: the Choices for our Energy Future, edited by Gerald Foley and Ariane van Buren; U.S. Energy Demand: Some Low Energy Futures; Coal: Bridge to the Future; Energy: Global Prospects 1985–2000. Bulletin of the Atomic Scientists, 36(9):43–47, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bastin:1986:LSF

[Bas86] Clinton Bastin. Letter: Spent-fuel storage: some background. Bulletin of the Atomic Scientists, 42(1):55–56, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BWA85].

Bateson:1946:PARa

[Bat46a] Gregory Bateson. The pattern of an armaments race: An anthropological approach — Part 1. Bulletin of the Atomic Scientists, 2(5–6):10–11, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bateson:1946:PARb

[Bat46b] Gregory Bateson. The pattern of an armaments race: An anthropological approach — Part 2 — an analysis of nationalism. Bulletin of the Atomic Scientists, 2(7–8):26–28, October 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bataille:1952:LEH

[Bat52] Edward F. Bataille. Letter to the Editor: How to lose friends. Bulletin of the Atomic Scientists, 8(3):96, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab52b].

Bateman:1961:MSD

[Bat61] Alan M. Bateman. Minerals: Supply and demand. Bulletin of the Atomic Scientists, 17(8):331–335, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bates:1969:IAS

[Bat69] Frederick L. Bates. The impact of automation on society. Bulletin of the Atomic Scientists, 25(6):4–6, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Batisse:1973:EPS

[Bat73] Michel Batisse. Environmental problems and the scientist. Bulletin of the Atomic Scientists, 29(2):15–21, February 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bates:1986:LIO

[Bat86] Don G. Bates. Letter: IPPNW’s official and unofficial roles. Bulletin of the Atomic Scientists, 42(8):54, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra86].

Batzel:1988:BLF

[Bat88] Roger E. Batzel. Batzel on Livermore’s fairness. Bulletin of the Atomic Scientists, 44(6):11, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Batchelor:1998:AA

[Bat98] Peter Batchelor. Arms and the ANC. Bulletin of the Atomic Scientists, 54(5):56–61, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bauer:1960:EDU

[Bau60] P. T. Bauer. Economic development of underdeveloped countries. Bulletin of the Atomic Scientists, 16(8):329–333, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bauer:1984:III

[Bau84] Etienne Bauer. Irresistible, irrational, indomitable military technology. Bulletin of the Atomic Scientists, 40(5):3–4, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Whe84a, Gre84b].

Bawden:1960:PD

[Baw60] F. C. Bawden. Plant diseases. Bulletin of the Atomic Scientists, 16(6):247–249, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bazelon:1983:TLJ

[Baz83] David L. Bazelon. Technology, litigation and justice. Bulletin of the Atomic Scientists, 39(9):10–11, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baker:1952:TOP

[BB52] John R. Baker and P. W. Bridgman. Two opinions on Professor Polanyi. Bulletin of the Atomic Scientists, 8(7):229, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Born:1963:BBS

[BB63] Hedwig Born and Max Born. Books: The Scientist Speculator: An Anthology of Partly-Baked Ideas, edited by I. J. Good. Bulletin of the Atomic Scientists, 19(5):30–32, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This is a strongly critical review of the book. The review begins: “The publication of Books: The Scientist Speculator: An Anthology of Partly-Baked Idea is nothing less than a crime against the ethical code, unwritten but vital, of the community of scientists.” See rebuttal in [Dys63a].

Brewer:1979:WGN

[BB79] Garry D. Brewer and Bruce G. Blair. War games and national security with a grain of SALT. Bulletin of the Atomic Scientists, 35 (6):18–26, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boardman:1987:LTL

[BB87] Donnell W. Boardman and Martha A. Bartter. Letters: Technostrategic language. Bulletin of the Atomic Scientists, 43(6):61–62, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coh87].

Black:2010:GTS

[BB10] Samuel Black and Yousaf Butt. The growing threat of space debris. Bulletin of the Atomic Scientists, 66(2):1–8, March 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/2/1.

Brill:2020:PPS

[BB20] Kenneth C. Brill and John H. Bernhard. Preventing the preventable: Strengthening international controls to thwart radiological terrorism. Bulletin of the Atomic Scientists, 76(4):206– 209, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1778371.

Byers:2022:CLS

[BB22] Michael Byers and Aaron Boley. Cis-lunar space and the security dilemma. Bulletin of the Atomic Scientists, 78(1):17–21, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.2014231.

Bunn:1998:RRN

[BBC^{+}98] Matthew Bunn, Oleg Bukharin, Jill Cetina, Kenneth Luongo, and Frank von Hippel. Retooling Russia’s nuclear cities. Bulletin of the Atomic Scientists, 54(5):44–50, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackett:1946:MAE

[BBD^{+}46] P. M. S. Blackett, M. Born, P. I. Dee, P. A. M. Dirac, N. Feather, E. A. Guggenheim, H. S. W. Massey, P. B. Moon, N. F. Mott, M. L. E. Oliphant, F. A. Paneth, R. E. Peierls, M. H. L. Pryce, F. E. Simon, Sir George Thompson, O. R. Frisch, and H. W. B. Skinner. Memo to the UN Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(12):6–8, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bopp:1957:DGN

[BBF^{+}57] Fritz Bopp, Max Born, Rudolf Fleischmann, Walter Gerlach, Otto Hahn, Otto Haxel, Werner Heisenberg, Hans Kopfermann, Max von Laue, Heinz Maier-Leibnitz, Josef Mattauch, Friedrich-Adolf Paneth, Wolfgang Paul, Wolfgang Riezler, Fritz Strassmann, Wilhelm Walcher, Carl Friedrich von Weizsäcker, and Karl Wirtz. Declaration of the German nuclear physicists: April 13, 1957. Bulletin of the Atomic Scientists, 13(6):228, June 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The declaration calls for the Federal Republic of Germany (‘West Germany’) to “renounce explicitly and voluntarily the possession of atomic weapons of any kind”, and states that none of the authors is ‘ready in any way to take part in the production, the tests, or the application of atomic weapons”. See followup [vW57].

Bacher:1982:UR

[BBF^{+}82] Robert F. Bacher, Hans A. Bethe, Adrian Fisher, Richard L. Garwin, Noel Gayler, John M. Lee, Norman F. Ramsey, Herbert Scoville, Jr., Paul C. Warnke, Victor F. Weisskopf, Jerome B. Wiesner, Robert R. Wilson, and Herbert F. York. USC recommendations. Bulletin of the Atomic Scientists, 38(6):5, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1985:OLD

[BBG^{+}85] Hans Bethe, Norris Bradbury, Richard Garwin, Spurgeon M. Keeny, Jr., Wolfgang Panofsky, George Rathjens, Herbert Scoville, Jr., and Paul Warnke. [Open letter to Dante Fascell, chairman, Committee on Foreign Affairs, U.S. House of Representatives]: Tests not necessary. Bulletin of the Atomic Scientists, 41(10):11, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1973:VCS

[BBH^{+}73] Hans A. Bethe, Robert F. Bacher, Philip Handler, Harvey Brooks, Bentley Glass, Alvin Weinberg, Frederick Seitz, Joseph Rotblat, Julian Schwinger, Cyril Stanley Smith, Walter Orr Roberts, Glenn T. Seaborg, John A. Simpson, Detlev W. Bronk, Harold C. Urey, Sewall Wright, Philip M. Morse, and I. I. Rabi. A voice of conscience is stilled. IV. messages. Bulletin of the Atomic Scientists, 29(6):8–12, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Eulogy for Eugene Rabinowitch (1901–1973), co-founder of the Bulletin.

Buzzard:1957:LAW

[BBHGA57] Sir Anthony Buzzard, P. M. S. Blackett, Denis Healey, and Richard Goold-Adams. On limiting atomic war. Bulletin of the Atomic Scientists, 13(6):216–222, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Condensation of a pamphlet for the Royal (British) Institute of International Affairs.

Bouveret:2013:NCN

[BBL13] Patrice Bouveret, Bruno Barrillot, and Dominique Lalanne. Nuclear chromosomes: The national security implications of a French nuclear exit. Bulletin of the Atomic Scientists, 69(1):11– 17, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/1/11.

Bentz:2014:AAD

[BBP14] Julie A. Bentz, Daniel J. Blumenthal, and Bradley A. Potter. It’s all about the data: Responding to international chemical, biological, radiological, and nuclear incidents. Bulletin of the Atomic Scientists, 70(4):57–68, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1950:FAHa

[BBSS50] Hans A. Bethe, Harrison S. Brown, Frederick Seitz, and Leo Szilard. The facts about the hydrogen bomb. Bulletin of the Atomic Scientists, 6(4):106–109, 126–127, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bin:2001:MDC

[BBZ01] Li Bin, Zhou Baogen, and Liu Zhiwei. Missile defense: China will have to respond. Bulletin of the Atomic Scientists, 57(6):25–28, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
6/25.

Burroughs:1996:NT

[BC96] John Burroughs and Jacqueline Cabasso. Nukes on trial. Bulletin of the Atomic Scientists, 52(2):41–45, March/April 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bradshaw:2016:BBG

[BC16] Michael Bradshaw and Richard Connolly. Barrels and bullets: The geostrategic significance of Russia’s oil and gas exports. Bulletin of the Atomic Scientists, 72(3):156–164, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burton:1981:CAM

[BCB81] John Burton, Barry Childers, and Glenn C. Buchan. Commentary: On ‘Anti-MAD Mythology’. Bulletin of the Atomic Scientists, 37(7):58–60, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Buc81].

Brown:1948:ESP

[BCH^{+}48] Harrison Brown, Karl Taylor Compton, Thorfin R. Hogness, Charles C. Lauritsen, Philip M. Morse, George B. Pegram, Harold C. Urey, and John C. Warner. Eight scientists protest Thomas Committee’s methods. Bulletin of the Atomic Scientists, 4(10):290, 320, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Telegram of 6 September 1948 to President Truman.

Bethe:1992:LLL

[BCMR92] Hans A. Bethe, George Cowan, Nicholas Metropolis, and Louis Rosen. Letter: Labs leap to own defense. Bulletin of the Atomic Scientists, 48(9):45–46, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blumstein:1974:LNM

[BCW74] Carl Blumstein, Robert D. Clear, and John C. Wheeler. Letter: A numerical mistake. Bulletin of the Atomic Scientists, 30(9): 3, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul74b, Mul74a].

Beale:1976:LE

[BD76] William T. Beale and Samuel H. Day, Jr. Letter: Exposed. Bulletin of the Atomic Scientists, 32(5):2, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bickham:1979:BRB

[BD79] Stephen H. Bickham and John Dowling. Book review: Nuclear War Films, edited by Jack G. Shaheen. Bulletin of the Atomic Scientists, 35(9):54–55, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Baltimore:1977:TSR

[BDE^{+}77] D. Baltimore, B. D. Davis, H. Eagle, J. T. Edsall, C. Grobstein, D. M. Horstmann, R. Hotchkiss, R. J. Huebner, A. Rich, W. A. Rosenblith, F. N. Ruddle, R. L. Sinsheimer, and E. L. Smith. A threat to scientific research. Bulletin of the Atomic Scientists, 33(10):9, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wri77]. The authors are all members of the U.S. National Academy of Sciences.

Berkhout:1992:PTS

[BDF^{+}92] Frans Berkhout, Anatoli Diakov, Harold Feiveson, Marvin Miller, and Frank von Hippel. Plutonium: True separation anxiety. Bulletin of the Atomic Scientists, 48(9):28–34, November 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Best:1962:LED

[BE62] Wallace H. Best and Burton I. Edelson. Letter to the Editor: A department of science. Bulletin of the Atomic Scientists, 18(7): 38, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spo62].

Beard:1947:CFC

[Bea47a] W. C. Beard. A. S. A. E. committee for foreign correspondence. Bulletin of the Atomic Scientists, 3(9):266, September 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beard:1947:ASA

[Bea47b] Walter C. Beard, Jr. Association of Scientists for Atomic Education news. Bulletin of the Atomic Scientists, 3(11):344, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beadle:1959:MVH

[Bea59] George W. Beadle. Molecules, viruses, and heredity. Bulletin of the Atomic Scientists, 15(9):354–360, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beadle:1964:NBN

[Bea64] George W. Beadle. The new biology and the nature of man. Bulletin of the Atomic Scientists, 20(3):13–17, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beals:1967:CCR

[Bea67] Ralph L. Beals. Cross-cultural research and government policy. Bulletin of the Atomic Scientists, 23(8):18–24, October 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bean:1986:UEC

[Bea86a] Kevin Bean. Untitled: [Economic conversion misses the point]. Bulletin of the Atomic Scientists, 42(6):48–50, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ada86a].

Beaton:1986:LHM

[Bea86b] Hugh R. Beaton. Letter: How many lives did the bombs justify? Bulletin of the Atomic Scientists, 42(9):56, November 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber86b].

Beard:1987:LRC

[Bea87] James E. Beard. Letter: Reactor conversion dangerous. Bulletin of the Atomic Scientists, 43(8):53–54, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hoe87].

Beach:1989:PCT

[Bea89] Sir Hugh Beach. Perspective: The case for the third zero. Bulletin of the Atomic Scientists, 45(10):14–15, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beckerley:1955:IGI

[Bec55] J. G. Beckerley. The impact of government information and security controls on competitive industry. Bulletin of the Atomic Scientists, 11(4):123–127, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Bechhoefer:1958:IAE

[Bec58] Bernhard C. Bechhoefer. The International Atomic Energy Agency. Bulletin of the Atomic Scientists, 14(4):147–150, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Becker:1962:LEP

[Bec62] Benjamin M. Becker. Letter to the Editor: Peace by force. Bulletin of the Atomic Scientists, 18(9):31–32, November 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot61].

Bechhoefer:1964:TBT

[Bec64] Berhard G. Bechhoefer. The test-ban treaty: Some further considerations. Bulletin of the Atomic Scientists, 20(5):25–27, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Becker:1971:MAC

[Bec71] Benjamin M. Becker. The myth of arms control and disarmament. Bulletin of the Atomic Scientists, 27(4):5–8, 45–48, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Becker:1975:UNA

[Bec75] Benjamin M. Becker. The United Nations after thirty years. Bulletin of the Atomic Scientists, 31(9):31–35, November 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sma76a].

Beckerly:1978:LSW

[Bec78] James C. Beckerly. Letter: Strauss’s wrath. Bulletin of the Atomic Scientists, 34(1):3, 5, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Becker:1979:CES

[Bec79] Benjamin M. Becker. Commentary: Even SALT II requires trust. Bulletin of the Atomic Scientists, 35(8):60, October 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bauer:1993:LMV

[BED93] S. H. Bauer, Elmer Eisner, and Sam Day. Letters: Mordechai Vanunu: Victim, villain — or both? Bulletin of the Atomic Scientists, 49(1):52–53, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day92].

Berry:1972:SRWa

[BEGJ72] R. Stephen Berry, Paul R. Ehrlich, Frances Gendin, and Sally Jacobsen. A special report: What happened at Stockholm. Bulletin of the Atomic Scientists, 28(7):15–16, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bridgman:1955:SAA

[BEI^{+}55] Percy W. Bridgman, Albert Einstein, Leopold Infeld, Hermann Joseph Muller, Cecil F. Powell, Bertrand Russell, and Hideki Yukawa. Scientists appeal for abolition of war. Bulletin of the Atomic Scientists, 11(7):236–237, September 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint in [RE63].

Beissinger:1982:SUP

[Bei82] Mark R. Beissinger. The Soviet Union: The power elite. Bulletin of the Atomic Scientists, 38(8):24–31, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beissinger:1984:BRB

[Bei84] Mark R. Beissinger. Book review: Industry Innovation in the Soviet Union, edited by R. Amann and J. M. Cooper. Bulletin of the Atomic Scientists, 40(5):42–44, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bell:1946:CMT

[Bel46] Bernard Iddings Bell. The common moral task. Bulletin of the Atomic Scientists of Chicago, 1(5):9, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Belehradek:1948:SSS

[Bél48] Jan Bélehrádek. A scientist speaks for the small nations. Bulletin of the Atomic Scientists, 4(2):62, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bell:1955:TC

[Bel55] Charles G. Bell. Titan chained. Bulletin of the Atomic Scientists, 11(5):180, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bell:1957:PVS

[Bel57] Hallen M. Bell. Part VII: Science and religion: Science and religion as social controls. Bulletin of the Atomic Scientists, 13(1): 32–34, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bellin:1972:CCG

[Bel72] J. S. Bellin. Communications: CBW and Geneva Protocol. Bulletin of the Atomic Scientists, 28(1):2, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bell:1993:BRS

[Bel93] John Bell. Book review: Smart kills: War in the Age of Intelligent Machines, by Manuel De Landa. Bulletin of the Atomic Scientists, 49(5):54, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Belyaninov:1994:NNB

[Bel94] Kirill Belyaninov. Nuclear nonsense, black-market bombs, and fissile flim-flam. Bulletin of the Atomic Scientists, 50(2):44–50, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bell:2002:ACB

[Bel02] Glenn Bell. Article commentary: Beryllium: Giving Kazakhstan the business. Bulletin of the Atomic Scientists, 58(6):17–18, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/58/
6/17.

Belcher:2020:TON

[Bel20a] Emma Belcher. Transforming our nuclear future with ridiculous ideas. Bulletin of the Atomic Scientists, 76(6):325–330, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1846420.

Bell:2020:WPC

[Bel20b] Alexandra Bell. What the presidential candidates should be asked about arms control and nonproliferation. Bulletin of the Atomic Scientists, 76(1):27–30, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1701801.

Benfey:1956:SCH

[Ben56] O. T. Benfey. The scientist’s conscience: Historical considerations. Bulletin of the Atomic Scientists, 12(5):177–178, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benton:1958:PC

[Ben58] William Benton. The pace of change. Bulletin of the Atomic Scientists, 14(5):162–163, 192, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bennett:1959:RAS

[Ben59] Rainey Bennett. The romance of art and science. Bulletin of the Atomic Scientists, 15(2):52–54, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benson:1961:LEB

[Ben61] R. Benson. Letter to the Editor: Bilateral peace corps. Bulletin of the Atomic Scientists, 17(4):120, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benarde:1969:PTC

[Ben69] Melvin A. Benarde. Pharmaceutical testing: Company view: Physician heal thyself. Bulletin of the Atomic Scientists, 25(6):17, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benedict:1971:EPN

[Ben71a] Manson Benedict. Electric power from nuclear fission. Bulletin of the Atomic Scientists, 27(7):8–16, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benn:1971:TPP

[Ben71b] Anthony Wedgwood Benn. I. Technical power and people: The impact of technology on the structure of government. Bulletin of the Atomic Scientists, 27(10):23–26, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benitez:1972:OOE

[Ben72] Helena Z. Benitez. Only one Earth: For whom? Bulletin of the Atomic Scientists, 28(7):55–56, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benoit:1976:CAS

[Ben76a] Emile Benoit. The coming age of shortages. Bulletin of the Atomic Scientists, 32(1):6–16, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Lau76, Gra76b].

Benoit:1976:DEE

[Ben76b] Emile Benoit. A dynamic equilibrium economy: the coming age of shortages. Bulletin of the Atomic Scientists, 32(2):47–55, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Hou76, Wes76, FD76].

Benoit:1976:FSS

[Ben76c] Emile Benoit. First steps to survival. Bulletin of the Atomic Scientists, 32(3):41–48, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benson:1978:LDW

[Ben78] Ruth Benson. Letter: DOE weapons. Bulletin of the Atomic Scientists, 34(3):4, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bennett:1979:MK

[Ben79a] Governor Robert F. Bennett. The MX in Kansas! Bulletin of the Atomic Scientists, 35(2):64, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Open letter to US President Jimmy Carter opposing a proposed MX missile site in Kansas.

Benson:1979:LIC

[Ben79b] Ruth Benson. Letter: an inspired cover. Bulletin of the Atomic Scientists, 35(3):86, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bennett:1983:CAN

[Ben83] Philip W. Bennett. Commentary: Adding to the numbers game. Bulletin of the Atomic Scientists, 39(4):45–46, April 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benda:1987:RL

[Ben87] Susan Benda. Reform legislation. Bulletin of the Atomic Scientists, 43(4):29, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Benn:2001:ASR

[Ben01a] Aluf Benn. Arms sales: The Russian–Iranian connection. Bulletin of the Atomic Scientists, 57(1):16–18, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/1/16.

Benn:2001:CWC

[Ben01b] Aluf Benn. Chemical weapons convention: Israel’s decision time. Bulletin of the Atomic Scientists, 57(2):22–24, March 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/2/22.

Benn:2001:ICP

[Ben01c] Aluf Benn. Israel: Censoring the past. Bulletin of the Atomic Scientists, 57(4):17–19, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/4/17.

Benn:2001:IQC

[Ben01d] Aluf Benn. Israel: Quiet cooperation. Bulletin of the Atomic Scientists, 57(6):15–17, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/15.

Benn:2002:MES

[Ben02] Aluf Benn. The Middle East: Slowing the missile race. Bulletin of the Atomic Scientists, 58(1):21–22, January/February 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/1/21.

Benedict:2007:SFI

[Ben07] Kennette Benedict. The sky is falling: an interview with John P. Holdren. Bulletin of the Atomic Scientists, 63(1):40–45, January/ February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
1/40.

Bernard:1946:STS

[Ber46] Chester I. Bernard. Security through the sacrifice of sovereignty. Bulletin of the Atomic Scientists, 2(7–8):30, October 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernal:1949:OSP

[Ber49] J. D. Bernal. Ought science to be planned? two opposing views: The case for collective research. Bulletin of the Atomic Scientists, 5(1):17–19, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pol49].

Berkner:1953:SMP

[Ber53a] Lloyd V. Berkner. Science and military power. Bulletin of the Atomic Scientists, 9(10):359–365, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Mar54b].

Berkner:1953:SNS

[Ber53b] Lloyd V. Berkner. Science and the national strength. Bulletin of the Atomic Scientists, 9(5):154–155, 180–181, June 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkner:1955:SE

[Ber55] Lloyd V. Berkner. Is secrecy effective? Bulletin of the Atomic Scientists, 11(2):62–63, 68, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkner:1956:PF

[Ber56] L. V. Berkner. Power of freedom. Bulletin of the Atomic Scientists, 12(5):174–176, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkner:1958:MSS

[Ber58] Lloyd V. Berkner. Man’s space satellites. Bulletin of the Atomic Scientists, 14(3):106–111, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkner:1963:CH

[Ber63] L. V. Berkner. The compelling horizon. Bulletin of the Atomic Scientists, 19(5):8–10, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berreman:1964:FIA

[Ber64] Gerald D. Berreman. Fear itself: An anthropologist’s view. Bulletin of the Atomic Scientists, 20(9):8–11, November 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1965:EWW

[Ber65] Henry T. Bernstein. East works with West. Bulletin of the Atomic Scientists, 21(1):37–39, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkner:1966:BBO

[Ber66a] L. V. Berkner. Books: The Organization of Research Establishments, by Sir John Cockcroft. Bulletin of the Atomic Scientists, 22(7):36–38, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berreman:1966:LER

[Ber66b] Gerald D. Berreman. Letter to the Editor: On the role of women. Bulletin of the Atomic Scientists, 22(9):26–28, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [DB67].

Berry:1970:CCC

[Ber70a] R. Stephen Berry. The chemistry and cost of contamination: Perspectives on polluted air — 1970. Bulletin of the Atomic Scientists, 26(4):2, 34–41, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berthoff:1970:MFL

[Ber70b] Ann E. Berthoff. Mumbling and fumbling — or lying? Bulletin of the Atomic Scientists, 26(8):49–50, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1971:OS

[Ber71] R. Stephen Berry. The option for survival. Bulletin of the Atomic Scientists, 27(5):22–27, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berenblum:1972:BAR

[Ber72a] I. Berenblum. Basic and applied research in a small country: The biomedical potential. Bulletin of the Atomic Scientists, 28 (8):29–33, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Berry:1972:CLG

[Ber72b] R. Stephen Berry. Comment: Limits to growth — two views: Reflections on The Limits to Growth. Bulletin of the Atomic Scientists, 28(9):25–27, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wat73b].

Berry:1972:RTE

[Ber72c] R. Stephen Berry. Recycling, thermodynamics and environmental thrift. Bulletin of the Atomic Scientists, 28(5):8–15, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [GB72].

Berry:1972:SRWb

[Ber72d] R. Stephen Berry. A special report: What happened at Stockholm: I. Only one world: an awakening. Bulletin of the Atomic Scientists, 28(7):17–20, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1973:EFY

[Ber73a] Donna Berry. ERTS — the first year in orbit. Bulletin of the Atomic Scientists, 29(6):41–43, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1973:FPD

[Ber73b] Donna Berry. Flood plain development: The Salt Creek disaster. Bulletin of the Atomic Scientists, 29(4):35–36, April 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1973:BBE

[Ber73c] R. Stephen Berry. Books: International Environmental Action — A Global Survey, by Thomas W. Wilson, Jr. Bulletin of the Atomic Scientists, 29(1):48–49, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berendzen:1975:BBC

[Ber75a] Richard Berendzen. Books: The Cosmic Connection, by Carl Sagan. Bulletin of the Atomic Scientists, 31(4):47–48, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1975:SWH

[Ber75b] Barton J. Bernstein. Shatterer of worlds: Hiroshima and Nagasaki. Bulletin of the Atomic Scientists, 31(10):12–22, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [San76a].

Berry:1975:CRS

[Ber75c] R. Stephen Berry. Crisis of resource scarcity: The transition to an energy-limited economy. Bulletin of the Atomic Scientists, 31(1): 31–36, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beres:1976:LNQ

[Ber76a] Louis R. Beres. Letter: Nuclear quills. Bulletin of the Atomic Scientists, 32(10):4–5, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [San76b].

Bernstein:1976:WWA

[Ber76b] Bargon J. Bernstein. The week we almost went to war. Bulletin of the Atomic Scientists, 32(2):12–21, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Hig76, You76c].

Berlau:1977:LWO

[Ber77a] A. J. Berlau. Letter: What oil shortage? Bulletin of the Atomic Scientists, 33(4):4, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1977:BSE

[Ber77b] R. Stephen Berry. Books: The search for an environmental ethic: When Values Conflict: Essays on Environmental Analysis, Discourse and Decision, edited by Laurence H. Tribe, Corinne S. Schelling, and John Voss. Bulletin of the Atomic Scientists, 33(3): 71–72, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1978:BBC

[Ber78a] Barton J. Bernstein. Books: Energy and Conflict: The Life and Times of Edward Teller, by Stanley Blumberg and Gwinn Owens. Bulletin of the Atomic Scientists, 34(5):51–53, May 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BO76, OB78].

Berry:1978:LFM

[Ber78b] Yale J. Berry. Letter: Free market oil. Bulletin of the Atomic Scientists, 34(8):64, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bergmann:1979:BRW

[Ber79] Peter G. Bergmann. Book reviews: Why I left Canada, by Leopold Infeld. Translated by Helen Infeld. Edited with introduction and notes by Lewis Pyenson. Foreword by Alfred Schild. Bulletin of the Atomic Scientists, 35(3):60–61, March 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

BRPF:1980:PEC

[Ber80] Bertrand Russell Peace Foundation. Perspective: The END campaign. Bulletin of the Atomic Scientists, 36(10):1, 54, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beres:1981:REH

[Ber81] Louis René Beres. 1980 Rabinowitch Essay on how to eliminate the threat of nuclear war: Steps toward a new planetary identity. Bulletin of the Atomic Scientists, 37(2):43–47, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beres:1982:CWP

[Ber82] Louis René Beres. Commentary: When politics distorts human rights. Bulletin of the Atomic Scientists, 38(2):43–44, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1983:CG

[Ber83] Michael Berry. Commentary: With gratitude. Bulletin of the Atomic Scientists, 39(2):51, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkowitz:1984:BRB

[Ber84a] Bruce D. Berkowitz. Book review: Strengthening Conventional Defense in Europe: Proposals for the 1980s, by The European Study Group. Bulletin of the Atomic Scientists, 40(1):39–40, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernardin:1984:PTT

[Ber84b] Joseph Cardinal Bernardin. Pacem in terris: twenty years later. Bulletin of the Atomic Scientists, 40(2):11–14, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1984:THB

[Ber84c] Barton J. Bernstein. Truman and the H-bomb. Bulletin of the Atomic Scientists, 40(3):12–18, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SB84].

Berry:1984:FL

[Ber84d] R. Stephen Berry. The federal laboratories. Bulletin of the Atomic Scientists, 40(3):21–25, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1985:RWP

[Ber85] Barton J. Bernstein. Radiological warfare: the path not taken. Bulletin of the Atomic Scientists, 41(7):44–49, August 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1986:CES

[Ber86a] Barton J. Bernstein. Chernobyl: the emerging story: Nuclear deception: the U.S. record. Bulletin of the Atomic Scientists, 42(7):40–43, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1986:PMU

[Ber86b] Barton J. Bernstein. A postwar myth: 500,000 U.S. lives saved. Bulletin of the Atomic Scientists, 42(6):38–40, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [MB86, Bea86b].

Bernstein:1986:LTN

[Ber86c] David Bernstein. Letter: Testing necessity. Bulletin of the Atomic Scientists, 42(6):61, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gaf86].

Berney:1987:LDL

[Ber87a] Arthur L. Berney. Letter: From deadly to lively competition. Bulletin of the Atomic Scientists, 43(2):59, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1987:BBR

[Ber87b] Barton J. Bernstein. Barton Bernstein responds: [Churchill’s anthrax bombs: a debate]. Bulletin of the Atomic Scientists, 43 (9):44–45, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1987:CSB

[Ber87c] Barton J. Bernstein. Churchill’s secret biological weapons. Bulletin of the Atomic Scientists, 43(1):46–50, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See debate [JL87, Ber87b].

Berend:1989:HBL

[Ber89a] Iván T. Berend. Hungary breaks loose. Bulletin of the Atomic Scientists, 45(5):27–31, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berkowitz:1989:ACC

[Ber89b] Bruce D. Berkowitz. Arms control criteria: Corollaries. Bulletin of the Atomic Scientists, 45(4):39–40, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berger:1993:ED

[Ber93a] Mikhail Berger. The economy: Disintegrating. Bulletin of the Atomic Scientists, 49(1):32–35, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bernstein:1993:SCT

[Ber93b] Barton J. Bernstein. Seizing the contested terrain of early nuclear history: Stimson, Conant, and their Allies explain the decision to use the atomic bomb. Diplomatic History, 17 (1):35–72, January 1993. ISSN 0145-2096 (print), 1467-7709

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1111/j.1467-7709.1993.tb00158.x.

Bernstein:1995:GOM

[Ber95] Barton J. Bernstein. Guest opinion: Misconceived patriotism. Bulletin of the Atomic Scientists, 51(3):4, May/June 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cam95b].

Bergeron:2001:WNO

[Ber01] Kenneth Bergeron. While no one was looking. Bulletin of the Atomic Scientists, 57(2):42–49, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/42.

Bernstein:2003:RAG

[Ber03] Barton J. Bernstein. Reconsidering the “Atomic General”: Leslie R. Groves. The Journal of Military History, 67(3):883–920, July 2003. ISSN 0899-3718 (print), 1543-7795 (electronic). URL http:

//www.jstor.org/stable/3397331.

Bergeron:2004:NWD

[Ber04a] Kenneth D. Bergeron. Nuclear weapons: The death of no-dual-use. Bulletin of the Atomic Scientists, 60(1):15–17, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/1/15.

Bernstein:2004:LTL

[Ber04b] Jeremy Bernstein. Larger than life. Bulletin of the Atomic Scientists, 60(2):35–41, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/2/35;   http://bos.sagepub.com/
content/60/2/35.abstract. Excerpted from [Ber04c].

Bernstein:2004:OPE

[Ber04c] Jeremy Bernstein. Oppenheimer: portrait of an enigma. Ivan R. Dee, Chicago, IL, USA, 2004. ISBN 1-56663-569-1. xi + 223 pp. LCCN QC16.O62 B43 2004.

Berrigan:2023:HMG

[Ber23] Frida Berrigan. How my Gen Z students learned to start worrying and dismantle the Bomb. Bulletin of the Atomic Scientists, 80(1):17–24, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2293569.

Bess:1985:LSS

[Bes85] Michael Bess. Leo Szilard: scientist, activist, visionary. Bulletin of the Atomic Scientists, 41(11):11–18, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Besmann:1994:LNS

[Bes94] Theodore M. Besmann. Letter: Nuclear success in France. Bulletin of the Atomic Scientists, 50(5):3, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot94d].

Brian:2002:WCW

[BES02] Danielle Brian, Lynn Eisenman, and Peter D. H. Stockton. The weapons complex: Who’s guarding the store? Bulletin of the Atomic Scientists, 58(1):48–55, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/48.

Bethe:1946:CAW

[Bet46] Hans A. Bethe. Can air or water be exploded? Bulletin of the Atomic Scientists, 1(7):2, 14, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1950:HBa

[Bet50] Hans A. Bethe. The hydrogen bomb. Bulletin of the Atomic Scientists, 6(4):99–104, 125, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bet63].

Bethe:1954:NAB

[Bet54] Hans A. Bethe. Negotiations and atomic bombs. Bulletin of the Atomic Scientists, 10(1):9–10, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1958:BBT

[Bet58] Hans A. Bethe. Books: Brighter Than a Thousand Suns, By Robert Jungk. Bulletin of the Atomic Scientists, 14(10):426–428, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1960:LEA

[Bet60] Hans A. Bethe. Letter to the Editor: Appraises Russian proposal. Bulletin of the Atomic Scientists, 16(5):144, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1962:AA

[Bet62a] Hans Bethe. [The Atomic Age]. Bulletin of the Atomic Scientists, 18(10):19, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1962:DS

[Bet62b] Hans A. Bethe. Disarmament and strategy. Bulletin of the Atomic Scientists, 18(7):14–22, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1963:AHB

[Bet63] Hans Bethe. The American hydrogen bomb. In Grodzins and Rabinowitch [GR63], pages 144–155. LCCN D842 .B78. Reprint of [Bet50].

Bettleheim:1965:BBA

[Bet65] Bruno Bettleheim. Books: Psychiatric Aspects of the Prevention of Nuclear War. Bulletin of the Atomic Scientists, 21(6):55–56, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1967:JRO

[Bet67] Hans A. Bethe. J. Robert Oppenheimer, 1904–1967. Bulletin of the Atomic Scientists, 23(8):3–6, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1969:ACA

[Bet69a] Hans A. Bethe. The ABM, China and the Arms Race. Bulletin of the Atomic Scientists, 25(5):41–44, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1969:MAM

[Bet69b] Hans A. Bethe. Missiles and anti-missiles … six views: Hard point vs. city defense. Bulletin of the Atomic Scientists, 25(6):25– 26, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Washington Scene: Missiles and Antimissiles.

Bethe:1970:DPP

[Bet70] Hans A. Bethe. Disarmament problems. Part 4 of “The Military Atom”. Bulletin of the Atomic Scientists, 26(6):99–102, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Betts:1972:CWC

[Bet72] Michele Betts. Communications: On Woman’s commandments. Bulletin of the Atomic Scientists, 28(5):2, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch72d, Sch72b].

Bethe:1975:NFC

[Bet75a] Hans A. Bethe. “No fundamental change in the situation” [nuclear reactor safety]. Bulletin of the Atomic Scientists, 31(7):40– 41, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1975:RSR

[Bet75b] Hans A. Bethe. Report of Steering Review Committee of the American Physical Society’s Study Group on Light Water Reactor Safety. Bulletin of the Atomic Scientists, 31(9):35–37, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1976:UC

[Bet76] Hans A. Bethe. Ultimate catastrophe? Bulletin of the Atomic Scientists, 32(6):36–37, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75b].

Bethe:1977:RNP

[Bet77a] Hans A. Bethe. Response: Nuclear power: a safe bet. Bulletin of the Atomic Scientists, 33(7):55, September 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet77b, Sel77, Fes77].

Bethe:1977:SVA

[Bet77b] Hans A. Bethe. Six views on atomic energy: The need for nuclear power. Bulletin of the Atomic Scientists, 33(3):59–63, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sel77, Bet77a].

Bethe:1981:EMS

[Bet81a] Hans A. Bethe. Editorial: Meaningless superiority. Bulletin of the Atomic Scientists, 37(8):1, 4, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1981:SCE

[Bet81b] Hans A. Bethe. Supernova collapse and explosion. Bulletin of the Atomic Scientists, 26(1):17, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1985:TI

[Bet85] Hans A. Bethe. The technological imperative. Bulletin of the Atomic Scientists, 41(7):34–36, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1986:UPA

[Bet86] Hans A. Bethe. U.S. panel assesses Chernobyl. Bulletin of the Atomic Scientists, 42(10):45–46, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bla87b].

Bethe:1989:CNA

[Bet89a] Hans A. Bethe. Chop down the nuclear arsenals. Bulletin of the Atomic Scientists, 45(2):11–15, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mol89].

Bethe:1989:LBM

[Bet89b] Hans A. Bethe. Letter: Better missiles. Bulletin of the Atomic Scientists, 45(7):44, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1990:NHS

[Bet90] Hans A. Bethe. Nuclear history: Sakharov’s H-bomb. Bulletin of the Atomic Scientists, 46(8):8–9, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://alsos.wlu.edu/information.aspx?id=2432.

Bethe:1993:LBG

[Bet93] Hans A. Bethe. Letter: Bethe on the German bomb program. Bulletin of the Atomic Scientists, 49(1):53–54, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See corrections [Ano93-31].

Bethe:1995:LCD

[Bet95] Hans A. Bethe. Letter: Cease and desist. Bulletin of the Atomic Scientists, 51(6):3, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Betancourt:1998:RNW

[Bet98] Alberto Betancourt. Report: Nuclear waste: Border skirmish. Bulletin of the Atomic Scientists, 54(3):14–16, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jag98].

Bethe:2020:CAW

[Bet20] H. A. Bethe. 1946: Can air or water be exploded? Bulletin of the Atomic Scientists, 76(6):341–343, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847466.

Beues:1949:BBP

[Beu49] Austin M. Beues. Books: No Place to Hide, by David Bradley. Bulletin of the Atomic Scientists, 5(4):128, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beveridge:1956:SBM

[Bev56] Lord [William Henry] Beveridge. [Speech before Milan conference, 4 August 1955]. Bulletin of the Atomic Scientists, 12(2): 43, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bevc:1986:LCW

[Bev86] Vladislav Bevc. Letter: Converting weapons labs. Bulletin of the Atomic Scientists, 42(6):64, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hug87].

Beyer:1946:LEP

[Bey46] Otto Beyer. Letter to the Editor: Personnel of the Atomic Energy Commission. Bulletin of the Atomic Scientists, 2(11–12):11, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beyea:1980:CDI

[Bey80a] Jan Beyea. Commentary: Dispute at Indian Point. Bulletin of the Atomic Scientists, 36(5):63–64, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BWH80].

Beyea:1980:EPR

[Bey80b] Jan Beyea. Emergency planning for reactor accidents. Bulletin of the Atomic Scientists, 36(10):40–45, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Beyea:1985:LCR

[Bey85] Jan Beyea. Letter: Critical responsibility. Bulletin of the Atomic Scientists, 41(11):63–64, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [CEE85].

Beyea:2012:SJP

[Bey12a] Jan Beyea. The scientific jigsaw puzzle: Fitting the pieces of the low-level radiation debate. Bulletin of the Atomic Scientists, 68 (3):13–28, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/3/13.

Beyea:2012:SIR

[Bey12b] Jan Beyea. Special issue on the risks of exposure to low-level radiation. Bulletin of the Atomic Scientists, 68(3):10–12, May/ June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/68/3/10.

Bezdek:1984:BH

[Bez84] Roger H. Bezdek. B-1: a history. Bulletin of the Atomic Scientists, 40(9):10–13, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berry:1973:ECA

[BF73] R. Stephen Berry and Margaret F. Fels. The energy cost of automobiles. Bulletin of the Atomic Scientists, 29(10):11–17, 58–60, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boureston:2004:SIA

[BF04] Jack Boureston and Charles D. Ferguson. Schooling Iran’s atom squad. Bulletin of the Atomic Scientists, 60(3):31–35, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/3/31.

Boureston:2005:ACK

[BF05a] Jack Boureston and Charles D. Ferguson. Article commentary: Keep your enemy closer. Bulletin of the Atomic Scientists, 61(6): 25, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/6/25.

Boureston:2005:LES

[BF05b] Jack Boureston and Charles D. Ferguson. Laser enrichment: Separation anxiety. Bulletin of the Atomic Scientists, 61(2): 14–18, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/2/14.

Bethe:1985:BSM

[BFG^{+}85] H. Bethe, R. Forsberg, R. Garwin, W. A. Higinbotham, P. Morrison, J. A. Simpson, C. Smith, and R. Wilson. The Bulletin and the scientists’ movement. Bulletin of the Atomic Scientists, 41 (11):19–31, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1949:SGN

[BFM^{+}49] Harrison Brown, James Franck, Joseph E. Mayer, Leo Szilard, and Harold C. Urey. Scientists give new warning. Bulletin of the Atomic Scientists, 5(10):264, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bok:1947:SRF

[BFW47] Bart J. Bok, Francis Friedman, and Victor Weiskopf. Security regulations in the field of nuclear research. Bulletin of the Atomic Scientists, 3(11):321–324, 344, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bezuidenhout:2014:WBA

[BG14] Louise Bezuidenhout and Chandre Gould. Winning the battle against emerging pathogens: a South African response. Bulletin of the Atomic Scientists, 70(4):10–13, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1960:LEW

[BGR60] Harrison Brown, Bentley Glass, and Eugene I. Rabinowitch. Letter to the Editor: What is Pugwash? Bulletin of the Atomic Scientists, 16(8):304, 345, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bludman:1961:BBN

[BGR61] S. A. Bludman, A. Garren, and A. H. Rosenfeld. Books: The Nation’s Safety and Arms Control, by Arthur T. Hadley; Arms Reduction: Program and Issues, edited by David H. Frisch; Strategy and Arms Control, by Thomas C. Schelling and Morton H. Halperin. Bulletin of the Atomic Scientists, 17(9):380–381, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bregman:1953:CMF

[BH53] Judith Bregman and I. Halpern. Cambridge meetings of FAS. Bulletin of the Atomic Scientists, 9(3):93, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Berman:1962:AEP

[BH62] William H. Berman and Lee M. Hydeman. Atomic energy policy review. Bulletin of the Atomic Scientists, 18(3):35–37, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1965:LEO

[BH65] J. C. Brown and Herbert Harvey. Letter to the Editor: Open responses to Levine. Bulletin of the Atomic Scientists, 21(1):31, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev64].

Blacker:1980:ETN

[BH80] Coit Dennis Blacker and Farooq Hussain. European theater nuclear forces. Bulletin of the Atomic Scientists, 36(8):32–37, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bhabha:1955:PUA

[Bha55] Homi J. Bhabha. The peaceful uses of atomic energy. Bulletin of the Atomic Scientists, 11(8):280–284, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bhargava:1970:CJ

[Bha70] P. M. Bhargava. Communications: On Jensenism. Bulletin of the Atomic Scientists, 26(9):40, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jen70].

Bhathal:1971:SGS

[Bha71] R. S. Bhathal. Science and government in Singapore. Bulletin of the Atomic Scientists, 27(1):20–21, 38, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bhaskar:1999:SC

[Bha99] Uday Bhaskar. Squaring the circle. Bulletin of the Atomic Scientists, 55(6):4, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/6/4.

Bhatia:2003:PA

[Bha03] Michael Bhatia. The peace allergy. Bulletin of the Atomic Scientists, 59(4):52–59, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/59/4/52.

Bowman:2011:MFT

[BHH11] Katherine Bowman, Kathryn Hughes, and Jo L. Husbands. Moving forward: Trends in science and technology and the future of the biological weapons convention. Bulletin of the Atomic Scientists, 67(3):16–25, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/67/3/16.

Buono:2021:TRU

[BHK^{+}21] Stephen Buono, Jake Hecla, Vladimir Kobezskii, Katie Mummah, and Julien de Troullioud de Lanversin. It’s time to reignite US–Russia cooperation in space. Nuclear power may hold the key. Bulletin of the Atomic Scientists, 77(4):203–206, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1941598.

Bowes:1991:LMF

[BHM91] Stanley A. Bowes, Daniel Hirsch, and William G. Mathews. Letters: More fallout from H-bomb controversy. Bulletin of the Atomic Scientists, 47(9):45–46, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90a] and comment [Han92c].

Biddle:1951:SF

[Bid51] Francis Biddle. Sickness of fear. Bulletin of the Atomic Scientists, 7(11):323–326, 336, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bienen:1968:BBC

[Bie68] Henry Bienen. Books: Peaceful Conflict: The Nonmilitary Use of the Military, by Edward Glick. Bulletin of the Atomic Scientists, 24(4):29–30, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bilder:1985:FTT

[Bil85] Richard B. Bilder. Formal treaties and tacit agreements: an exchange. Bulletin of the Atomic Scientists, 41(4):51–53, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See rejoinder [Har85c].

Bilderback:1986:PM

[Bil86] Loy Bilderback. Population: Migration. Bulletin of the Atomic Scientists, 42(4):33–37, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bingham:1954:PPP

[Bin54] Jonathan B. Bingham. Partisan politics and Point Four. Bulletin of the Atomic Scientists, 10(3):84–86, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Binder:1967:MEC

[Bin67] Leonard Binder. The Middle East crisis: A trial balance. Bulletin of the Atomic Scientists, 23(7):2–7, 33–37, September 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Binder:1979:SRR

[Bin79] Leonard Binder. Special report: Revolution in Iran: red, white, blue or black. Bulletin of the Atomic Scientists, 35(1):48–54, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Binder:1984:AME

[Bin84] Leonard Binder. Agenda 1984: The Middle East. Bulletin of the Atomic Scientists, 40(1):12–14, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bin:2004:CWC

[Bin04] Li Bin. China: Weighing the costs. Bulletin of the Atomic Scientists, 60(2):21–23, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/2/21.

Bin:2010:AVN

[Bin10] Li Bin. An alternative view to North Korea’s bomb acquisition. Bulletin of the Atomic Scientists, 66(3):38–43, May 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/3/38.

Bin:2015:ZRT

[Bin15] Li Bin. Is “zero” the right target for disarmament?: a Chinese response. Bulletin of the Atomic Scientists, 71(1):98–101, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bin:2018:CAT

[Bin18] Li Bin. China’s attitudes toward missile defense and its limitation. Bulletin of the Atomic Scientists, 74(4):243–247, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1486601.

BEIRC:1973:BRE

[Bio73] Biological Effects of Ionizing Radiation Committee. The BEIR report: Effects on populations of exposure to low levels of ionizing radiation. Bulletin of the Atomic Scientists, 29(3):47–49, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Birnbaum:1976:LSA

[Bir76] Saul Birnbaum. Letter: Socialism the answer. Bulletin of the Atomic Scientists, 32(8):2, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com76b].

Birnbaum:1985:OEE

[Bir85] Norman Birnbaum. 1985 outlook: Europeanization of Europe. Bulletin of the Atomic Scientists, 41(1):20–22, January 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Birnbaum:1987:BRB

[Bir87] Norman Birnbaum. Book review: Watershed in Europe: Dismantling the East–West Military Confrontation, by Jonathan Dean. Bulletin of the Atomic Scientists, 43(10):42–43, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Birkholz:1989:GPT

[Bir89] Mario Birkholz. German physicists: 1,700 tests are enough. Bulletin of the Atomic Scientists, 45(5):37, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bissio:2014:BEI

[Bis14] Roberto Bissio. Biomass energy and the implications for climate and food: The Uruguayan response. Bulletin of the Atomic Scientists, 70(1):9–11, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/70/1/9.full.pdf+html.

Bixler:1949:SMH

[Bix49] Julius S. Bixler. Science, materialism, and the human spirit: The affinities of science with morality and religion. Bulletin of the Atomic Scientists, 5(6–7):194–196, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bachar:1988:LBW

[BJCP88] John Bachar, Jack R. Jennings, Antonie K. Churg, and Sheldon C. Plotkin. Letter: Behind the Woodruff disclosure. Bulletin of the Atomic Scientists, 44(7):60, September 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu88].

Broyles:1968:LEW

[BJH68] Arthur A. Broyles, Stanley L. Jaki, and William Cornelius Hall. Letters to the Editor: On the Wigner–Feld exchange. Bulletin of the Atomic Scientists, 24(5):45–46, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel67b, WF68].

Benard:1981:CSN

[BK81] Cheryl Benard and Zalmay Khalilzad. Commentary: Still ‘No quick fix’. Bulletin of the Atomic Scientists, 37(10):59, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [KB80].

Burr:2003:NNP

[BK03] William Burr and Jeffrey Kimball. Nixon’s nuclear ploy. Bulletin of the Atomic Scientists, 59(1):28, January/February 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/1/28.

Butler:1960:RCS

[BKSG60] Allan M. Butler, Frederick G. Keyes, and Albert Szent-Gyorgy. Reader’s comment: Sea disposal of atomic wastes. Bulletin of the Atomic Scientists, 16(4):141, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burney:2013:GSA

[BKV13] Jennifer A. Burney, Charles F. Kennel, and David G. Victor. Getting serious about the new realities of global climate change. Bulletin of the Atomic Scientists, 69(4):49–57, July 2013. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/69/4/49.

Ball:1979:FAS

[BL79] Nicole Ball and Milton Leitenberg. The foreign arms sales of the Carter administration. Bulletin of the Atomic Scientists, 35(2): 31–36, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Broda:1981:CBJ

[BL81] Englebert Broda and Robert Jay Lifton. Commentary: On behalf of Joliot. Bulletin of the Atomic Scientists, 37(3):61, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lif80].

Bertram:1983:CNG

[BL83a] Christoph Bertram and Milton Leitenberg. Commentary: The numbers game. Bulletin of the Atomic Scientists, 39(1):49–50, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei82b].

Bethe:1983:EFR

[BL83b] Hans A. Bethe and Franklin A. Long. Editorial: The freeze referendum: what next? Bulletin of the Atomic Scientists, 39(2): 2–3, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Black:1983:CFI

[BL83c] Edwin F. Black and Leona M. Libby. Commercial food irradiation. Bulletin of the Atomic Scientists, 39(6):48–50, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackett:1946:AEA

[Bla46] P. M. S. Blackett. Atomic energy and The UNO Atomic Energy Commission. Bulletin of the Atomic Scientists, 1(8):14–15, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackett:1957:TWA

[Bla57] P. M. S. Blackett. Technology and world advancement. Bulletin of the Atomic Scientists, 13(9):323–326, 342, November 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackett:1961:NSO

[Bla61a] P. M. S. Blackett. New science or old technology? Bulletin of the Atomic Scientists, 17(2):53–54, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blagonravov:1961:BFC

[Bla61b] A. A. Blagonravov. Breaking the feedback cycle. Bulletin of the Atomic Scientists, 17(4):144–146, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackett:1962:UM

[Bla62] P. M. S. Blackett. The University’s mission. Bulletin of the Atomic Scientists, 18(5):14–18, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blagonravov:1967:CSR

[Bla67] Anatoly Blagonravov. Comments: Space research justified by needs of science. Bulletin of the Atomic Scientists, 23(8):40–42, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bor68].

Blackett:1969:CSC

[Bla69a] P. M. S. Blackett. Chicago: The social control of science and its applications: The ever-widening gap. Bulletin of the Atomic Scientists, 25(5):23–25, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blagonravov:1969:SRJ

[Bla69b] A. A. Blagonravov. A Soviet reply to joint flights. Bulletin of the Atomic Scientists, 25(7):15, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Black:1977:OS

[Bla77] Max Black. The objectivity of science. Bulletin of the Atomic Scientists, 33(2):55–60, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Black:1982:BRB

[Bla82a] Jan Knippers Black. Book review: El Salvador: Background to the Crisis. Bulletin of the Atomic Scientists, 38(8):46–48, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackaby:1982:WA

[Bla82b] Frank Blackaby. World arsenals 1982. Bulletin of the Atomic Scientists, 38(6):21–26, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blacker:1984:BRB

[Bla84] Coit D. Blacker. Book reviews: Beyond the Hotline: Controlling a Nuclear Crisis, by William Langer Ury and Richard Smoke; Avoiding Inadvertent War: Crisis Management, edited by Hilliard Roderick and Ulla Magnusson; Between Peace and War: The Nature of International Crisis, by Richard Ned Lebow. Bulletin of the Atomic Scientists, 40(7):43–45, August/ September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blacker:1985:BRB

[Bla85a] Coit D. Blacker. Book review: Survival Is Not Enough: Soviet Realities and America’s Future, by Richard Pipes. Bulletin of the Atomic Scientists, 41(4):55–56, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cas85a].

Blacker:1985:LUS

[Bla85b] Coit D. Blacker. Lessons from U.S.–Soviet summits. Bulletin of the Atomic Scientists, 41(10):14–17, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blacker:1987:BRB

[Bla87a] Coit Blacker. Book review: Military Objectives in Soviet Foreign Policy, by Michael MccGwire. Bulletin of the Atomic Scientists, 43(6):55–57, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackith:1987:LRS

[Bla87b] Robert Blackith. Letter: RBMK safety. Bulletin of the Atomic Scientists, 43(3):52, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet86].

Blanck:1988:LCR

[Bla88] George Blanck. Letter: Conflagration without representation. Bulletin of the Atomic Scientists, 44(10):49, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackaby:1997:GOT

[Bla97] Frank Blackaby. Guest opinion: Time for a peasants’ revolt. Bulletin of the Atomic Scientists, 53(6):4, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mia98].

Blanton:2002:UGZ

[Bla02] Thomas Blanton. Umbra Gamma Zarf. Bulletin of the Atomic Scientists, 58(1):62–63, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/62.

Blair:2007:PR

[Bla07] Bruce G. Blair. Primed and ready. Bulletin of the Atomic Scientists, 63(1):33–37, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/33.

Blair:2011:FFA

[Bla11] Charles P. Blair. Fatwas for fission: Assessing the terrorist threat to Pakistan’s nuclear assets. Bulletin of the Atomic Scientists, 67 (6):19–33, November/December 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/67/6/19.

Blair:2020:LCP

[Bla20] Bruce G. Blair. Loose cannons: The President and US nuclear posture. Bulletin of the Atomic Scientists, 76(1):14–26, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1701279.

Bless:1954:LEM

[Ble54] A. A. Bless. Letter to the Editor: Madame Joliot and the American Chemical Society. Bulletin of the Atomic Scientists, 10(6): 213, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blechman:1987:BRB

[Ble87] Barry M. Blechman. Book review: Nuclear crisis management: A Dangerous Illusion, by Richard Ned Lebow. Bulletin of the Atomic Scientists, 43(7):55–56, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blechman:1988:MRM

[Ble88] Barry M. Blechman. A minimal reduction of a major risk. Bulletin of the Atomic Scientists, 44(3):44–46, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blechman:1989:ACC

[Ble89] Barry M. Blechman. Arms control criteria: Cost reduction dubious. Bulletin of the Atomic Scientists, 45(4):38–39, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BPG89].

Bliss:1970:BEW

[Bli70] L. C. Bliss. A biologist explains: Why we must plan now to protect the Arctic. Bulletin of the Atomic Scientists, 26(8):34– 38, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bliss:1976:WJB

[Bli76] Raymond W. Bliss. Why not just build the house right in the first place? Bulletin of the Atomic Scientists, 32(3):32–40, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bloomfield:1963:POS

[Blo63] Lincoln P. Bloomfield. The politics of outer space. Bulletin of the Atomic Scientists, 19(5):12–14, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bloomfield:1966:LEH

[Blo66] Lincoln P. Bloomfield. Letter to the Editor: How to keep the peace. Bulletin of the Atomic Scientists, 22(8):38–39, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RP66].

Bloomfield:1971:ANI

[Blo71] Lincoln P. Bloomfield. After neo-isolationism, what? Bulletin of the Atomic Scientists, 27(4):9–13, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bloomfield:1985:NCH

[Blo85] Lincoln P. Bloomfield. Nuclear crisis and human frailty. Bulletin of the Atomic Scientists, 41(9):26–30, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bloomfield:1995:LET

[Blo95] Janet Bloomfield. Letter: In England, they’re dumps. Bulletin of the Atomic Scientists, 51(3):69, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jag95].

Bloom:2005:MDS

[Blo05] Mia Bloom. Mother. daughter. sister. bomber. Bulletin of the Atomic Scientists, 61(6):54–62, November/December 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/6/54.

Blumenstein:1984:CO

[Blu84a] Carl R. Blumenstein. Commentary: Overkill. Bulletin of the Atomic Scientists, 40(10):35, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ACH84].

Blumenstein:1984:CSE

[Blu84b] Carl R. Blumenstein. Commentary: Sakharov and East–West relations. Bulletin of the Atomic Scientists, 40(9):54, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gor85].

Blum:1988:WSL

[Blu88] Deborah Blum. Weird science: Livermore’s X-ray laser flap. Bulletin of the Atomic Scientists, 44(6):7–13, July/August 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BJCP88, DeW88, Cun88].

Blume:2023:CDA

[Blu23] Lesley M. M. Blume. Collateral damage: American civilian survivors of the 1945 Trinity test. Bulletin of the Atomic Scientists, 79(4):232–237, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2223078.

Bruce:1993:LWJ

[BM93] Frank R. Bruce and David Marples. Letters: Who are the jurors at Chernobyl? Bulletin of the Atomic Scientists, 49(9):53–54, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar93a].

Bruce:1999:EWF

[BM99] Maxwell Bruce and Tom Milne, editors. Ending war: the force of reason: essays in honour of Joseph Rotblat, NL, FRS.

Macmillan Press, Houndmills, Basingstoke, Hampshire, UK, 1999. ISBN 0-333-76070-0 (Macmillan Press: hardcover), 0-333-77482-5 (Macmillan Press: paperback), 0-312-22570-9 (St. Martin’s Press: hardcover). xxiii + 179 pp. LCCN JZ5665 .E53 1999.

URL http://www.loc.gov/catdir/bios/hol057/99026118.
html;   http://www.loc.gov/catdir/description/hol056/
99026118.html;   http://www.loc.gov/catdir/toc/hol053/
99026118.html.

Bengelsdorf:2002:ICE

[BM02] Harold D. Bengelsdorf and Fred McGoldrick. International custody of excess plutonium. Bulletin of the Atomic Scientists, 58 (2):31–35, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/2/31.

Bunn:2008:NRN

[BM08] Matthew Bunn and Martin B. Malin. A nuclear revival needs new cooperation. Bulletin of the Atomic Scientists, 64(4):60, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/64/4/60.

Braden:2022:FCS

[BM22] Zoe Braden and Allison Macfarlane. The final countdown to site selection for Canada’s nuclear waste geologic repository. Bulletin of the Atomic Scientists, 79(1):22–27, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2155007.

BrittodeCastro:1989:BNS

[BMPd89] Antonio Rubens Britto de Castro, Norberto Majlis, Luiz Pinguelli Rosa, and Fernando de Souza Barros. Brazil’s nuclear shakeup: military still in control. Bulletin of the Atomic Scientists, 45(4): 22–25, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blackburn:2018:SVE

[BMPH18] Jim Blackburn, Henk Mooiweer, Megan Parks, and Annie Hutson. The Soil Value Exchange: Unlocking nature’s value via the market. Bulletin of the Atomic Scientists, 74(3):162–169, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1461974.

Bonney:1967:CRD

[BN67] Walter T. Bonney and H. L. Nieburg. Comments: R and D and the contract state. Bulletin of the Atomic Scientists, 23(2):28– 29, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nie66a].

Blumberg:1976:ECL

[BO76] Stanley A. Blumberg and Gwinn Owens. Energy and conflict: the life and times of Edward Teller. Putnam, New York, NY, USA, 1976. ISBN 0-399-11551-X. xvii + 492 + 4 pp. LCCN QC16.T37 B58 1976. US$12.95. URL http://alsos.wlu.edu/

information.aspx?id=367.

Bang:1979:CDW

[BØ79] Jens Bang and P. L. Ølgaard. Commentary from Denmark: a warning. Bulletin of the Atomic Scientists, 35(5):49–50, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

BD:1991:RNE

[Boa91] Board of Directors. Report: New editor appointed. Bulletin of the Atomic Scientists, 47(4):10, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

BD:1998:NMM

[Boa98] Board of Directors. Nine minutes to midnight. Bulletin of the Atomic Scientists, 54(5):4, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

BoD:2007:IMM

[Boa07] Board of Directors. It is 5 minutes to midnight. Bulletin of the Atomic Scientists, 63(1):66–71, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/1/66.

Bobrow:1965:RAN

[Bob65] Davis B. Bobrow. Realism about nuclear spread. Bulletin of the Atomic Scientists, 21(10):20–22, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boe:1952:LEC

[Boe52] R. E. Boe. Letter to the Editor: [Christian democracy]. Bulletin of the Atomic Scientists, 8(1):33, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boeri:1957:AFI

[Boe57a] Enzo Boeri. Academic freedom in Italian universities. Bulletin of the Atomic Scientists, 13(1):24–26, 38, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Don57] and reply [Boe57b].

Boeri:1957:LEA

[Boe57b] Enzo Boeri. Letter to the Editor: Academic freedom in Italian universities. Bulletin of the Atomic Scientists, 13(4):152, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Boe57a, Don57].

Boelentin:1976:LRB

[Boe76] Wanda Boelentin. Letter: The real Bucky Fuller. Bulletin of the Atomic Scientists, 32(1):3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ehr75].

Bohr:1950:OW

[Boh50] Niels Bohr. For an open world. Bulletin of the Atomic Scientists, 6(7):213–217, 219, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bohr:1955:G

[Boh55] Niels Bohr. On Geneva. Bulletin of the Atomic Scientists, 11(8): 284, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bok:1948:UPS

[Bok48] Bart J. Bok. UNESCO and the physical sciences. Bulletin of the Atomic Scientists, 4(11):343–347, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bok:1949:FSU

[Bok49] Bart J. Bok. Freedom of science and the Universal Declaration of Human Rights. Bulletin of the Atomic Scientists, 5(8–9):211– 217, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bolling:1951:PD

[Bol51] Richard Bolling. Politics in dispersal. Bulletin of the Atomic Scientists, 7(9):278–279, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bolton:1963:RVC

[Bol63] Charles D. Bolton. Resolving vicious circles in international affairs. Bulletin of the Atomic Scientists, 19(4):17–20, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bolt:1988:PAL

[Bol88a] Richard Bolt. Plutonium for all: leaks in global safeguards. Bulletin of the Atomic Scientists, 44(10):14–19, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Hir89b, FB89].

Bolt:1988:RBL

[Bol88b] Richard Bolt. Reading between the lines [nuclear safeguards]. Bulletin of the Atomic Scientists, 44(10):18, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bolt:1988:TMI

[Bol88c] Richard Bolt. Terms and methods: IAEA safeguards. Bulletin of the Atomic Scientists, 44(10):15, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bond:1949:MCC

[Bon49] Horatio Bond. Military and civil confusion about civil defense. Bulletin of the Atomic Scientists, 5(11):295–297, 314, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bond:1950:FAC

[Bon50] Horatio Bond. Fire aspects of civil defense. Bulletin of the Atomic Scientists, 6(8–9):246–248, August/September 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Bond:1951:DAF

[Bon51] Horatio Bond. Defense against fire effects of atomic bombs. Bulletin of the Atomic Scientists, 7(11):337–340, 352, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bond:1953:FSH

[Bon53a] Horatio Bond. Fire storm hazards. Bulletin of the Atomic Scientists, 9(7):277–279, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bontecou:1953:PES

[Bon53b] Eleanor Bontecou. President Eisenhower’s “Security” program. Bulletin of the Atomic Scientists, 9(6):215–217, 220, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bondy:1957:LVM

[Bon57] François Bondy. L’atome — à la vie, à la mort. (French) [The atom — in life and in death]. Bulletin of the Atomic Scientists, 13(9):346, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from a review of Ralph Lapp’s Atoms and People and Robert Jungk’s Heller als Tausend Sonnen from Preuves, March 1957, pp. 87–88.

Bondy:1958:LVMa

[Bon58a] François Bondy. L’atome — à la vie, à la mort. (French) [The atom — in life and in death]. Bulletin of the Atomic Scientists, 14 (1):66, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bondy:1958:LVMb

[Bon58b] François Bondy. L’atome — à la vie, à la mort. (French) [The atom — in life and in death]. Bulletin of the Atomic Scientists, 14 (2):97, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bonfiglio:1972:CPP

[Bon72] Albert Bonfiglio. Communications: Physics and politics in Latin America. Bulletin of the Atomic Scientists, 28(2):4, 48, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fal70].

Bonheim:1983:CA

[Bon83] Ralph Bonheim. Commentary: In agreement. Bulletin of the Atomic Scientists, 39(1):51, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eps82b].

Bonheim:1990:LCN

[Bon90] Ralph Bonheim. Letter: Code name tells all. Bulletin of the Atomic Scientists, 46(4):53, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack90k].

Boorse:1946:TIS

[Boo46] Henry A. Boorse. Two international scientific meetings in England. Bulletin of the Atomic Scientists, 2(5–6):5–6, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Booth:2022:SBP

[Boo22] Mary S. Booth. “Sustainable” biomass: a paper tiger when it comes to reducing carbon emissions. Bulletin of the Atomic Scientists, 78(3):139–147, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2062938.

Borst:1946:KM

[Bor46] Lyle Borst. Kansas meetings. Bulletin of the Atomic Scientists, 1 (10):14, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borgese:1948:WMC

[Bor48] Elizabeth Mann Borgese. Why a maximalist constitution? Bulletin of the Atomic Scientists, 4(7):199–204, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com48a].

Borgese:1952:BHS

[Bor52a] Elizabeth Mann Borgese. Books: History without statesmen: Homo Sapiens, by Hans Thirring. Bulletin of the Atomic Scientists, 8(6):198–201, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borgnis:1952:SOE

[Bor52b] F. E. Borgnis. Some other European experiences. Bulletin of the Atomic Scientists, 8(7):248, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Borst:1953:NWT

[Bor53] Lyle B. Borst. Nevada weapons test [March 17, 1953]. Bulletin of the Atomic Scientists, 9(3):73–75, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sal53].

Born:1957:MA

[Bor57a] Max Born. Man and the atom. Bulletin of the Atomic Scientists, 13(6):186–194, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bor63]. See note in [BR57].

Born:1957:MFB

[Bor57b] Max Born. [My former belief in the superiority of science]. Bulletin of the Atomic Scientists, 13(1):31, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1958:CRP

[Bor58a] Max Born. The concept of reality in physics. Bulletin of the Atomic Scientists, 14(8):313–321, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1958:ES

[Bor58b] Max Born. Europe and science. Bulletin of the Atomic Scientists, 14(2):73–79, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1958:RCS

[Bor58c] Max Born. Reader’s comment: Space flight for what? Bulletin of the Atomic Scientists, 14(6):234, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1960:PP

[Bor60] Max Born. Physics and politics. Bulletin of the Atomic Scientists, 16(6):194–200, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1963:MA

[Bor63] Max Born. Man and the atom. In Grodzins and Rabinowitch [GR63], pages 590–601. LCCN D842 .B78. Reprint of [Bor57a].

Born:1964:WLH

[Bor64] Max Born. What is left for hope for? Bulletin of the Atomic Scientists, 20(4):2–5, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1965:RMBa

[Bor65a] Max Born. Recollections of Max Born. I. How I became a physicist. Bulletin of the Atomic Scientists, 21(7):3–6, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1965:RMBb

[Bor65b] Max Born. Recollections of Max Born. II. What I did as a physicist. Bulletin of the Atomic Scientists, 21(8):9–13, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Born:1965:RMBc

[Bor65c] Max Born. Recollections of Max Born. III. Reflections. Bulletin of the Atomic Scientists, 21(9):3–6, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wil67a].

Born:1966:BES

[Bor66] Max Born. Blessings and evils of space travel. Bulletin of the Atomic Scientists, 22(8):12–14, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translated by Charles W. Scherr from the German original, published in Universitas and in Neue Deutsche Hefte, 106, 113 (1965), as well as in the author’s 1965 book of essays, Von der Verantwortung Naturwissenschaftlers.

Born:1968:CCM

[Bor68] Max Born. Current comments: From Max Born. Bulletin of the Atomic Scientists, 24(2):29, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ure67, Bla67].

Borisov:1969:CWC

[Bor69] P. W. Borisov. Can we control the Arctic climate? Bulletin of the Atomic Scientists, 25(3):43–48, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borgstrom:1970:DCH

[Bor70] Goerg A. Borgstrom. The dual challenge of health and hunger: a global crisis. Bulletin of the Atomic Scientists, 26(8):42–46, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borlaug:1971:GRB

[Bor71] Norman E. Borlaug. The green revolution: For bread and peace. Bulletin of the Atomic Scientists, 27(6):6–9, 42–48, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint of the author’s 1971 Nobel Peace Prize address.

Borlaug:1973:CFC

[Bor73] Norman E. Borlaug. Civilization’s future a call for international granaries. Bulletin of the Atomic Scientists, 29(8):7–15, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borawski:1982:CAC

[Bor82] John Borawski. Commentary: Arms control in Europe: building confidence. Bulletin of the Atomic Scientists, 38(1):45–46, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borawski:1985:LLP

[Bor85] John Borawski. Letter: Lack of progress at Stockholm. Bulletin of the Atomic Scientists, 41(8):64, September 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar85b].

Borawski:1986:AS

[Bor86a] John Borawski. Accord at Stockholm. Bulletin of the Atomic Scientists, 42(10):34–36, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borawski:1986:PST

[Bor86b] John Borawski. Progress in Stockholm talks. Bulletin of the Atomic Scientists, 42(2):40–42, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borawski:1987:USM

[Bor87] John Borawski. U.S.–Soviet move toward risk reduction. Bulletin of the Atomic Scientists, 43(6):16–18, July/August 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Borenstein:1989:MLN

[Bor89] Nathaniel S. Borenstein. My life as a NATO collaborator. Bulletin of the Atomic Scientists, 45(3):14–20, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SB89].

Borrie:2014:IPH

[Bor14] John Borrie. Improving UN planning for a humanitarian response to a nuclear detonation. Bulletin of the Atomic Scientists, 70 (5):74–85, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boskey:1951:ACP

[Bos51] Bennett Boskey. AEC contract policy and operations: Ninth semiannual report of the USAEC: Part two. Bulletin of the Atomic Scientists, 7(7–8):227–228, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Part I [Tur51].

Bosma:1978:LEN

[Bos78] John T. Bosma. Letter: Elegant nonsense. Bulletin of the Atomic Scientists, 34(2):2–3, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar77].

Boserup:1981:DD

[Bos81] Anders Boserup. Deterrence and defense. Bulletin of the Atomic Scientists, 37(10):11–13, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boserup:1988:WUH

[Bos88] Anders Boserup. A way to undermine hostility. Bulletin of the Atomic Scientists, 44(7):16–19, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bossie:1998:LHC

[Bos98] Robert Bossie, SCJ. Letter: The human cost. Bulletin of the Atomic Scientists, 54(4):68, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LC98b].

Bosco:2005:WAB

[Bos05] David Bosco. The world according to Bolton. Bulletin of the Atomic Scientists, 61(4):24–31, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/4/24.

Bosco:2006:BLR

[Bos06a] David Bosco. Bosnia: The Long road home. Bulletin of the Atomic Scientists, 62(2):19–20, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/2/19.

Bosco:2006:IP

[Bos06b] David Bosco. An indefinable problem. Bulletin of the Atomic Scientists, 62(1):44–51, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/1/44.

Bothwell:1961:ACI

[Bot61] Frank E. Bothwell. Arms control: The initiative is ours. Bulletin of the Atomic Scientists, 17(4):125–127, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bothwell:1969:IO

[Bot69] Frank E. Bothwell. Is the ICBM obsolete? Bulletin of the Atomic Scientists, 25(8):21–22, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1961:BBE

[Bou61] K. E. Boulding. Books: The Economics of Defense in the Nuclear Age, by Charles J. Hitch and Roland N. McKean. Bulletin of the Atomic Scientists, 17(3):115–116, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1962:ACW

[Bou62] Kenneth E. Boulding. After civilization what? Bulletin of the Atomic Scientists, 18(8):2–6, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1965:FRP

[Bou65] Kenneth E. Boulding. Forum: Reflections on protest. Bulletin of the Atomic Scientists, 21(8):18–20, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1966:NPP

[Bou66] Kenneth E. Boulding. Notes on the politics of peace. Bulletin of the Atomic Scientists, 22(7):30–32, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1970:SR

[Bou70] Kenneth E. Boulding. The scientific revelation. Bulletin of the Atomic Scientists, 26(7):13–18, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alb71].

Boulding:1974:WWW

[Bou74] Kenneth E. Boulding. What went wrong, if anything, since Copernicus? Bulletin of the Atomic Scientists, 30(1):17–23, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boucher:1977:LNO

[Bou77] Philippe Boucher. Letter: The nuclear opposition. Bulletin of the Atomic Scientists, 33(10):2, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1980:PLP

[Bou80] Kenneth E. Boulding. Perspective: So little for peace. Bulletin of the Atomic Scientists, 36(3):1, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1982:EP

[Bou82] Elise Boulding. Education for peace. Bulletin of the Atomic Scientists, 38(6):59–62, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulding:1985:NO

[Bou85] Elise Boulding. Nongovernmental organizations. Bulletin of the Atomic Scientists, 41(7):94–96, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boutwell:1987:BRB

[Bou87] Jeffrey Boutwell. Book review: Weaponry in Space: The Dilemma of Security, by Yevgeni Velikhov, Roald Sagdeyev, and Andrei Kokoshin. Bulletin of the Atomic Scientists, 43(7):54–55, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulden:1997:HS

[Bou97] Laurie H. Boulden. Harvest season. Bulletin of the Atomic Scientists, 53(5):30–34, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boulden:1998:PD

[Bou98a] Laurie H. Boulden. De-politicizing demining. Bulletin of the Atomic Scientists, 54(4):17–19, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bourgholtzer:1998:SVS

[Bou98b] Frank Bourgholtzer. Book review: Spy vs. spy: Battleground Berlin: CIA vs KGB in the Cold War, by David E. Murphy, Sergei A. Kondrashev, and George Bailey. Bulletin of the Atomic Scientists, 54(2):62–63, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boutwell:1999:WWB

[Bou99] Jeffrey Boutwell. The Wild West Bank. Bulletin of the Atomic Scientists, 55(1):41–44, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/41.

Bourgholtzer:2000:CP

[Bou00] Frank Bourgholtzer. Containment, part 1. Bulletin of the Atomic Scientists, 56(6):68–69, November/December 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/6/68.

Bourgholtzer:2002:BRB

[Bou02] Frank Bourgholtzer. Book review: Blame Truman: Another Such Victory: President Truman and the Cold War, 1945–1953, by Arnold A. Offner, Stanford University, 2002. 626 pages. Bulletin of the Atomic Scientists, 58(6):101–102, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/6/101.

Boutwell:2011:MTW

[Bou11] Jeffrey Boutwell. Moving toward a WMD-free Middle East. Bulletin of the Atomic Scientists, 67(1):74–75, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/1/74.

Bowles:1957:NAF

[Bow57] Chester Bowles. A new approach to foreign aid. Bulletin of the Atomic Scientists, 13(2):42–47, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bowie:1989:ACC

[Bow89] Robert R. Bowie. Arms control criteria: A promising future. Bulletin of the Atomic Scientists, 45(4):41, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boureston:2010:EIN

[BOW10] Jack Boureston and Tanya Ogilvie-White. Expanding the IAEA’s nuclear security mandate. Bulletin of the Atomic Scientists, 66 (5):55–64, September/October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/66/5/55.

Boyer:1984:AAA

[Boy84] Paul Boyer. From activism to apathy: America and the nuclear issue, 1963–1980. Bulletin of the Atomic Scientists, 40(7):14– 23, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WB85b].

Boyer:1985:SSB

[Boy85] Paul Boyer. Social scientists and the bomb. Bulletin of the Atomic Scientists, 41(9):31–36, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boyer:1986:HVS

[Boy86a] Paul Boyer. A historical view of scare tactics. Bulletin of the Atomic Scientists, 42(1):17–19, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [AC86, Cal86a].

Boyer:1986:LBE

[Boy86b] Paul Boyer. Letter: By the bomb’s early light. Bulletin of the Atomic Scientists, 42(6):62–63, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [For86].

Boyer:1987:BRB

[Boy87] Paul Boyer. Book review: The Iron Curtain: Churchill, America, and the Origins of the Cold War, by Fraser J. Harbutt. Bulletin of the Atomic Scientists, 43(4):47–48, 50, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pat87].

Boyer:1989:ARS

[Boy89a] Paul Boyer. Arms race as sitcom plot. Bulletin of the Atomic Scientists, 45(5):6–8, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mac89].

Boyer:1989:BRB

[Boy89b] Paul Boyer. Book review: Retreat from Doomsday: The Obsolescence of Major War, by John Mueller. Bulletin of the Atomic Scientists, 45(8):38–39, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boyer:1991:BRH

[Boy91] Paul Boyer. Book review: HUAC’s heyday: The Culture of the Cold War, by Stephen J. Whitfield. Bulletin of the Atomic Scientists, 47(5):40–41, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Boyer:1995:BRR

[Boy95] Paul D. Boyer. Book review: The rise and fall of Project Chariot: The Firecracker Boys, by Daniel T. O’Neill. Bulletin of the Atomic Scientists, 51(3):62–65, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://www.highbeam.com/doc/1G1-16888670.html.

Boyer:2000:LAL

[Boy00] Paul Boyer. Lofty aims and low intrigues. Bulletin of the Atomic Scientists, 56(5):59–61, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/59.

Boyer:2005:NTB

[Boy05] Paul Boyer. The nostalgia trap: Book review: Atomic Culture: How We Learned to Stop Worrying and Love the Bomb, edited by Scott C. Zeman and Michael A. Amundson. Bulletin of the Atomic Scientists, 61(1):67–69, January/February 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/1/67;   http://
www.highbeam.com/doc/1G1-126748359.html. See [ZA04].

Boyd:2010:UTW

[Boy10] Dallas Boyd. Unconventional thinking: Why conventional disarmament must precede nuclear abolition. Bulletin of the Atomic Scientists, 66(2):46–60, March 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/66/2/46.

Bolef:1984:PC

[BP84] Dan I. Bolef and George Palmer. On the plutonium connection. Bulletin of the Atomic Scientists, 40(6):56, June/July 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barus:1985:LSP

[BP85] Carl Barus and Gary Pereira. Letter: Sakharov’s plight. Bulletin of the Atomic Scientists, 41(4):60, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vL85].

Binkley:1990:PPS

[BP90] Cameron Binkley and William C. Potter. Perspective: Plutonium shipments safer by sub. Bulletin of the Atomic Scientists, 46 (10):14–15, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pat88, TN90].

Bukharin:1995:PWG

[BP95] Oleg Bukharin and William Potter. “Potatoes were guarded better”. Bulletin of the Atomic Scientists, 51(3):46–50, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bennett:1989:LAC

[BPG89] Paul C. Bennett, David J. Polewka, and Gene Garan. Letters: Arms control assessed. Bulletin of the Atomic Scientists, 45(6): 51–52, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ble89].

Bauer:1952:LED

[BPR52] Norman Bauer, Leonard T. Pockman, and Eugene Rabinowitch. Letter to the Editor: [drift toward one-sidedness in the Bulletin]. Bulletin of the Atomic Scientists, 8(3):96, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bauer:1976:NEF

[BPTB76a] Etienne Bauer, Louis Puiseux, and Pierre-Frédéric Ténière-Buchot. Nuclear energy: a fateful choice for France. Bulletin of the Atomic Scientists, 32(1):37–41, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zim76], rebuttal [Lap76] and response [BPTB76b].

Bauer:1976:SFC

[BPTB76b] Etienne Bauer, Louis Puiseux, and Pierre-Frédéric Ténière-Buchot. Still a fateful choice. Bulletin of the Atomic Scientists, 32(10):45, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BPTB76a, Lap76].

Born:1957:LEW

[BR57] Max Born and Eugene Rabinowitch. Letter to the Editor: We’re sorry. Bulletin of the Atomic Scientists, 13(7):273, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bauer:1978:P

[BR78] Etienne Bauer and Georges Ripka. Is this peace? Bulletin of the Atomic Scientists, 34(6):15–16, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1979:CPU

[BR79] Peter G. Brown and William L. Ranken. Commentary: A proxy for the unborn. Bulletin of the Atomic Scientists, 35(9):64–65, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bella:1985:LND

[BR85] David A. Bella and James D. Rise. Letter: New deployment mode for Midgetman. Bulletin of the Atomic Scientists, 41(5):54–55, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor84b].

Blechman:1990:CCB

[BR90] Barry M. Blechman and George Rathjens. Counterpoint: On cutting the budget in half. Bulletin of the Atomic Scientists, 46 (3):16–17, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burr:1997:CP

[BR97] William Burr and Jeffrey T. Richelson. A Chinese puzzle. Bulletin of the Atomic Scientists, 53(4):42–47, July/August 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Byravan:2015:SLR

[BR15] Sujatha Byravan and Sudhir Chella Rajan. Sea level rise and climate change exiles: a possible solution. Bulletin of the Atomic Scientists, 71(2):21–28, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bradley:1947:GBC

[Bra47] General Omar N. Bradley. General Bradley calls for international control — even at the expense of national sovereignty. Bulletin of the Atomic Scientists, 3(2):34, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From an address delivered at the “Report from the World” meeting on January 10, 1947.

Bradley:1949:GBO

[Bra49] Omar N. Bradley. General Bradley outlines American defense plans. Bulletin of the Atomic Scientists, 5(12):344, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bradley:1950:OUD

[Bra50] W. H. Bradley. Occurrence of uranium deposits. Bulletin of the Atomic Scientists, 6(5):149–152, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bradbury:1954:AL

[Bra54] Norris E. Bradbury. The Los Alamos Laboratory. Bulletin of the Atomic Scientists, 10(9):358–359, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bradley:1958:WSD

[Bra58] Omar N. Bradley. What others say: 1. A decent respect for human intelligence. Bulletin of the Atomic Scientists, 14(2):66, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Braverman:1974:CSS

[Bra74] Max Braverman. Communications: Steady state population. Bulletin of the Atomic Scientists, 30(3):3, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul74b].

Brandt:1984:A

[Bra84] Willy Brandt. Alternatives. Bulletin of the Atomic Scientists, 40(6):3–4, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brandt:1986:AHa

[Bra86a] Willy Brandt. Arms and hunger. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-52127-X (paperback). 208 pp. LCCN HC59.7 .B67713 1987. Translation by Anthea Bell of Der organisierte Wahnsinn.

Brandt:1986:AHb

[Bra86b] Willy Brandt. Arms and hunger. Pantheon Books, New York, NY, USA, 1986. ISBN 0-394-55446-9. 208 pp. LCCN HC59.7 .B67713 1986. Translation by Anthea Bell of Der organisierte Wahnsinn.

Braden:1987:BRB

[Bra87] Kathleen Braden. Book review: Arms and Hunger, by Willy Brandt. Bulletin of the Atomic Scientists, 43(2):54–56, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brac:1989:BRB

[Bra89] Timothy V. Brac. Book review: The Diplomacy of Biological Disarmament, by Nicholas A. Sims. Bulletin of the Atomic Scientists, 45(6):45–46, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bradley:1995:LCW

[Bra95] James H. Bradley. Letter: The Cold War — made in the U.S.A. Bulletin of the Atomic Scientists, 51(3):68–69, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo94e].

Bratt:2002:CNS

[Bra02] Duane Bratt. Canada’s nuclear schizophrenia. Bulletin of the Atomic Scientists, 58(2):44–50, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/44.

Bradford:2009:NRM

[Bra09] Peter A. Bradford. The nuclear renaissance meets economic reality. Bulletin of the Atomic Scientists, 65(6):60–64, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/6/60.

Braasch:2013:CCS

[Bra13a] Gary Braasch. Climate change: Is seeing believing? Bulletin of the Atomic Scientists, 69(6):33–41, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/6/33.

Bradford:2013:HCU

[Bra13b] Peter A. Bradford. How to close the US nuclear industry: Do nothing. Bulletin of the Atomic Scientists, 69(2):12–21, March/ April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/69/2/12.

Bradford:2017:WTS

[Bra17] Peter A. Bradford. Wasting time: Subsidies, operating reactors, and melting ice. Bulletin of the Atomic Scientists, 73 (1):13–16, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2016.1264207.

Branch:2016:CEC

[BRB16] Glenn Branch, Josh Rosenau, and Minda Berbeco. Climate education in the classroom: cloudy with a chance of confusion. Bulletin of the Atomic Scientists, 72(2):89–96, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brennan:1959:WOS

[Bre59] D. G. Brennan. Why outer space control? Bulletin of the Atomic Scientists, 15(5):198–202, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brennan:1965:BBC

[Bre65] D. G. Brennan. Books: Strategy and Conscience, by Anatol Rapoport. Bulletin of the Atomic Scientists, 21(10):25–30, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Rap65b].

Brennan:1967:NTM

[Bre67] D. G. Brennan. New thoughts on missile defense. Bulletin of the Atomic Scientists, 23(6):10–15, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brennan:1968:CCU

[Bre68] D. G. Brennan. Current comments: Uncertainty is not the issue. Bulletin of the Atomic Scientists, 24(3):33–34, March 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hib68].

Brennan:1969:CCS

[Bre69a] D. G. Brennan. Current comments: A start of strategic stabilization. Bulletin of the Atomic Scientists, 25(1):35–36, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brenner:1969:FND

[Bre69b] Michael J. Brenner. France’s new defence strategy and the Atlantic puzzle. Bulletin of the Atomic Scientists, 25(9):4–7, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brenner:1972:DDE

[Bre72] Michael J. Brenner. Decoupling, disengagement and European defense. Bulletin of the Atomic Scientists, 28(2):38–42, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brenner:1973:BBD

[Bre73] Michael J. Brenner. Books: In Defense of People, by Richard Neuhaus. Bulletin of the Atomic Scientists, 29(2):44–45, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brenner:1974:BBS

[Bre74] Michael Brenner. Books: Law of the Sea: Oceanic Resources, by Erin Bain Jones. Bulletin of the Atomic Scientists, 30(3):47–48, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Breslauer:1982:SAA

[Bre82] George W. Breslauer. Specious arguments for the arms race. Bulletin of the Atomic Scientists, 38(8):38–42, October 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brennan:1984:CNU

[Bre84] Robert X. Brennan. Commentary: On ‘Normalizing the unthinkable’. Bulletin of the Atomic Scientists, 40(6):56, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pea84].

Breen:1985:LVG

[Bre85] Susan Breen. Letter: Voters’ guide. Bulletin of the Atomic Scientists, 41(2):56–, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Breyman:1987:BRB

[Bre87] Steven J. Breyman. Book review: The SDI challenge to Europe, by Ivo H. Daalder. Bulletin of the Atomic Scientists, 43(9):49– 50, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Breyman:1993:BRP

[Bre93] Steve Breyman. Book review: Post-Nuclear Times: Between Fear & Hope: A Decade of Peace Activism Compiled from Nuclear Times Magazine: 1982 to Present, edited by Sonia Shah. Bulletin of the Atomic Scientists, 49(5):52–53, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brenner:2013:EWS

[Bre13] Joel F. Brenner. Eyes wide shut: The growing threat of cyber attacks on industrial control systems. Bulletin of the Atomic Scientists, 69(5):15–20, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/5/15.

Bridgman:1948:SSR

[Bri48a] Percy W. Bridgman. Scientists and social responsibility. Bulletin of the Atomic Scientists, 4(3):69–72, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ure48b, Rab48a, Gol48, DuB48, Ogb48, Rab48h].

Brimble:1948:ET

[Bri48b] L. J. F. Brimble. The exposition of the truth. Bulletin of the Atomic Scientists, 4(5):141–144, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bridgman:1949:SMH

[Bri49] Percy W. Bridgman. Science, materialism, and the human spirit: The scientist’s commitment. Bulletin of the Atomic Scientists, 5(6–7):192–193, 196, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

BASA:1950:CSP

[Bri50] British Atomic Scientists’ Association. The Civil Service purge in Britain. Bulletin of the Atomic Scientists, 6(6):185, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

BASA:1957:SH

[Bri57] British Atomic Scientists’ Association. Strontium hazards. Bulletin of the Atomic Scientists, 13(6):202–203, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bridgman:1958:SI

[Bri58] Percy W. Bridgman. Society and the individual. Bulletin of the Atomic Scientists, 14(10):413–416, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See note [Rab59p].

Brickman:1980:PSS

[Bri80a] Ronald Brickman. The price of success: science and technology in Europe. Bulletin of the Atomic Scientists, 36(7):29–34, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Briggs:1980:BRB

[Bri80b] Asa Briggs. Book review: Science in Victorian Manchester, by R. H. Kargon. Bulletin of the Atomic Scientists, 36(6):52, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brians:1986:BRB

[Bri86] Paul Brians. Book reviews: Galápagos, by Kurt Vonnegut; Fiskadoro, by Denis Johnson; The Nuclear Age, by Tim O’Brien; Countdown to Midnight: Twelve Great Stories about Nuclear War, edited by H. Bruce Franklin; Beyond Armageddon: Twenty-One Sermons to the Dead, edited by Walter M. Miller, Jr. and Martin H. Greenberg; Afterwar, edited by Janet Morris. Bulletin of the Atomic Scientists, 42(3):50–53, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brians:1987:NHA

[Bri87] Paul Brians. Nuclear Holocausts: Atomic War in Fiction, 1895– 1984. Kent State University Press, Kent, OH, USA, 1987. ISBN 0-87338-335-4. xi + 398 pp. LCCN PN3352.N83; Z5917.N83 B74 1987.

Brians:1988:NFC

[Bri88] Paul Brians. Nuclear fiction for children. Bulletin of the Atomic Scientists, 44(6):24–27, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brians:1989:BRB

[Bri89] Paul Brians. Book review: War Stars: The Superweapon and the American Imagination, by H. Bruce Franklin. Bulletin of the Atomic Scientists, 45(4):48, 50–51, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brin:2005:SC

[Bri05] David Brin. Smile for the camera. Bulletin of the Atomic Scientists, 61(4):64–67, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/4/64.

Brodie:1947:CAN

[Bro47a] Bernard Brodie. A critique of Army and Navy thinking on the atomic bomb. Bulletin of the Atomic Scientists, 3(8):207–210, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brodie:1947:NDT

[Bro47b] Bernard Brodie. Navy Department thinking on the atomic bomb. Bulletin of the Atomic Scientists, 3(7):177–180, 198–199, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1947:BEA

[Bro47c] H. Brown. Books: Explaining atoms to the people: The Atomic Story, by John W. Campbell; Meet the Atoms, by O. R. Frisch; Explaining the Atom by Selig Hecht. Bulletin of the Atomic Scientists, 3(6):163, 166, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1947:BER

[Bro47d] Harrison Brown. The beginning of the end: A review. Bulletin of the Atomic Scientists, 3(3):99, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1947:LEI

[Bro47e] Harrison Brown. Letter to the Editor: Important omission in account of world government. Bulletin of the Atomic Scientists, 3(7):197, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1947:WGM

[Bro47f] Harrison Brown. The world government movement in the United States. Bulletin of the Atomic Scientists, 3(6):156–157, 168, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brodie:1948:ABP

[Bro48] Bernard Brodie. The atom bomb as policy maker. Bulletin of the Atomic Scientists, 4(12):377–383, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brodie:1949:AER

[Bro49] Bernard Brodie. The atomic explosion in Russia: What is the outlook now? Bulletin of the Atomic Scientists, 5(10):268, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Browne:1951:R

[Bro51] Sir William Browne. Reply. Bulletin of the Atomic Scientists, 7(11):325, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1954:BBC

[Bro54] Harrison Brown. Books: The Challenge of Our Times, edited by Farrington Daniels and Thomas M. Smith. Bulletin of the Atomic Scientists, 10(7):297, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brode:1955:BRBa

[Bro55a] Bernice Brode. Book review: The Man in the Thick Lead Suit, by Daniel Lang. Bulletin of the Atomic Scientists, 11(7):267– 268, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brode:1955:BRBb

[Bro55b] Bernice Brode. Book review: Uranium Country, by Kathleen Brüyn. Bulletin of the Atomic Scientists, 11(10):376, 379, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1955:LSM

[Bro55c] Ralph S. Brown, Jr. Loyalty–security measures and employment opportunities. Bulletin of the Atomic Scientists, 11(4):113–117, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty. See erratum [Ano55j].

Brondsted:1956:WPE

[Bro56a] H. V. Brondsted. Warning and promise of experimental embryology. Bulletin of the Atomic Scientists, 12(3):66–75, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bronowski:1956:RRS

[Bro56b] J. Bronowski. The real responsibilities of the scientist. Bulletin of the Atomic Scientists, 12(1):10–13, 20, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1956:PIS

[Bro56c] Harrison Brown. Part IV: Science and world economy: A changed and changing world. Bulletin of the Atomic Scientists, 12(8):303– 307, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brode:1957:BBE

[Bro57a] Bernice Brode. Books: Atomic Energy is Your Business. Today’s Key to Tomorrow’s Profits, by Arnold Kramish and Eugene M. Zuckert . Bulletin of the Atomic Scientists, 13(7):272, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brodie:1957:BBW

[Bro57b] Bernard Brodie. Books: Atomic Weapons and East–West Relations, by P. M. S. Blackett. Bulletin of the Atomic Scientists, 13(3):107–108, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brodie:1957:NWC

[Bro57c] Bernard Brodie. Nuclear weapons and changing strategic outlooks. Bulletin of the Atomic Scientists, 13(2):56–61, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1957:HVW

[Bro57d] Harrison Brown. How vulnerable are we? Bulletin of the Atomic Scientists, 13(9):318–322, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1957:RWR

[Bro57e] Ralph S. Brown, Jr. Regression in the Wright Report. Bulletin of the Atomic Scientists, 13(7):253–256, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1958:TPC

[Bro58] Harrison Brown. Third Pugwash Conference: Science, technology, and world development. Bulletin of the Atomic Scientists, 14 (10):409–412, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brode:1960:BBH

[Bro60a] Bernice Brode. Books: From Hiroshima to the Moon: Chronicles of Life in the Atomic Age, by Daniel Lang. Bulletin of the Atomic Scientists, 16(1):29, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Broido:1960:MFF

[Bro60b] A. Broido. Mass fires following nuclear attack. Bulletin of the Atomic Scientists, 16(10):409–413, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1960:DIU

[Bro60c] Harold Brown. Detection and identification of underground nuclear explosions. Bulletin of the Atomic Scientists, 16(3):89–92, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1960:DNE

[Bro60d] Harold Brown. The detection of nuclear explosions. Bulletin of the Atomic Scientists, 16(2):49, 80, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1961:BBP

[Bro61] Seyom Brown. Books: A Proposal to Neutralize Nuclear Weapons, by Thornton Read. Bulletin of the Atomic Scientists, 17(4):162, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brower:1962:NRN

[Bro62a] Michael Brower. News & reports: Nuclear strategy of the Kennedy administration. Bulletin of the Atomic Scientists, 18 (8):34–41, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1962:GET

[Bro62b] Harrison Brown. Guest editorial: The twentieth year. Bulletin of the Atomic Scientists, 18(10):2–3, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Broido:1963:SFE

[Bro63a] A. Broido. Surviving fire effects of nuclear detonations. Bulletin of the Atomic Scientists, 19(3):20–23, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1963:UND

[Bro63b] Seyom Brown. A universal nuclear deterrent force. Bulletin of the Atomic Scientists, 19(4):20–24, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Broda:1964:WER

[Bro64] Engelbart Broda. When exchange is not really exchange. Bulletin of the Atomic Scientists, 20(9):23–25, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bronston:1966:LEP

[Bro66] William G. Bronston. Letter to the Editor: The physician and Vietnam. Bulletin of the Atomic Scientists, 22(9):24, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1967:PAE

[Bro67a] Allan H. Brown. The post-Apollo era —- decisions facing Nasa. Bulletin of the Atomic Scientists, 23(4):11–16, April 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1967:PEW

[Bro67b] Harrison Brown. The political-economic web: Crisis in development. Bulletin of the Atomic Scientists, 23(10):2–7, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bronowski:1968:SHD

[Bro68] J. Bronowski. Science as a humanistic discipline. Bulletin of the Atomic Scientists, 24(8):33–38, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ner69, Bro69].

Bronowski:1969:CBR

[Bro69] J. Bronowski. Communications: Bronowski replies. Bulletin of the Atomic Scientists, 25(5):49, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ner69, Bro68].

BNL:1970:CFM

[Bro70] Brookhaven National Laboratory. Communications: Fallout and Marshallese. Bulletin of the Atomic Scientists, 26(3):45, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Oft71].

Brown:1971:STD

[Bro71] Harrison Brown. Science, technology and the developing countries. Bulletin of the Atomic Scientists, 27(6):10–14, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1972:ARS

[Bro72] Harrison Brown. An American renaissance? A scientist’s view of what America can be. Bulletin of the Atomic Scientists, 28 (6):7–12, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1974:ANT

[Bro74] Bruce J. Brown. Atmospheric nuclear testing: A survey of medical statistics in Australia. Bulletin of the Atomic Scientists, 30 (2):13–15, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1975:FHS

[Bro75] Harrison Brown. The fissioning of human society. Bulletin of the Atomic Scientists, 31(1):6–7, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brockdorff:1976:LNI

[Bro76a] Cay Brockdorff. Letter: Nuclear insurance. Bulletin of the Atomic Scientists, 32(7):2, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Broda:1976:SPP

[Bro76b] E. Broda. Solar power: the photochemical alternative. Bulletin of the Atomic Scientists, 32(3):49–52, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brody:1979:CRC

[Bro79a] T. P. Brody. Commentary: A reverse chain reaction. Bulletin of the Atomic Scientists, 35(10):58, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brothers:1979:BRB

[Bro79b] Dwight Brothers. Book review: Politics and Markets: The World’s Political–Economic Systems, by Charles E. Lindblom. Bulletin of the Atomic Scientists, 35(7):44–46, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1979:EBE

[Bro79c] Harrison Brown. An early brief encounter [Albert Einstein and Andrei Gromyko]. Bulletin of the Atomic Scientists, 35(3):17–19, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Brown:1979:LHL

[Bro79d] Harrison Brown. Learning how to live in a technological society: I. The crisis of affluence. Bulletin of the Atomic Scientists, 35 (7):13–17, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1979:LLT

[Bro79e] Harrison Brown. Learning to live in a technological society. II. Toward a world community. Bulletin of the Atomic Scientists, 35(8):23–27, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brooks:1980:ESC

[Bro80a] Harvey Brooks. Energy: a summary of the CONAES report. Bulletin of the Atomic Scientists, 36(2):23–30, February 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1980:CGG

[Bro80b] Lester R. Brown. Commentary: “Going, going, gone” — at $3 a gallon! Bulletin of the Atomic Scientists, 36(3):58, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1981:BRB

[Bro81] Lester R. Brown. Book review: Unavailable At Any Price. Bulletin of the Atomic Scientists, 37(6):46–47, June/July 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Browder:1982:SAF

[Bro82] Felix E. Browder. Science and the American future. Bulletin of the Atomic Scientists, 38(9):26–30, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1984:AHF

[Bro84a] Lester R. Brown. Automobiles: a hazy future. Bulletin of the Atomic Scientists, 40(5):20–24, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1984:CCT

[Bro84b] Lowell Brown. Commentary: Contradiction in terms. Bulletin of the Atomic Scientists, 40(5):63–64, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:GPI

[Bro85a] Archie Brown. Gorbachev’s policy innovations. Bulletin of the Atomic Scientists, 41(10):18–22, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano86t].

Brown:1985:EBN

[Bro85b] Harrison Brown. From the Editors: A bombardment of nonsense. Bulletin of the Atomic Scientists, 41(6):3, June/July 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:EFF

[Bro85c] Harrison Brown. From the Editors: Fallout and falsehoods. Bulletin of the Atomic Scientists, 41(4):3, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:ELP

[Bro85d] Harrison Brown. From the Editors; linking past and future. Bulletin of the Atomic Scientists, 41(7):4–7, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:ENG

[Bro85e] Harrison Brown. From the Editors: Numbers game won’t work. Bulletin of the Atomic Scientists, 41(3):3–4, March 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:ETSc

[Bro85f] Harrison Brown. From the Editors: Test Soviet intentions. Bulletin of the Atomic Scientists, 41(11):3, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:ETSa

[Bro85g] Harrison Brown. From the Editors: Toward a strategy for peace. Bulletin of the Atomic Scientists, 41(9):3, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cox86].

Brown:1985:ETSb

[Bro85h] Harrison Brown. From the Editors: Toward a strategy for peace — II. Bulletin of the Atomic Scientists, 41(10):3–4, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1985:SWO

[Bro85i] Harrison Brown. Star Wars once funny, now frightening. Bulletin of the Atomic Scientists, 41(5):3, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [ATL85].

Brown:1986:BPT

[Bro86a] Harrison Brown. Breaking the public trust. Bulletin of the Atomic Scientists, 42(7):3, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1986:EGP

[Bro86b] Harrison Brown. From the Editors: The Gorbachev proposals. Bulletin of the Atomic Scientists, 42(3):3, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1986:ESN

[Bro86c] Harrison Brown. From the Editors: The summit and the new year. Bulletin of the Atomic Scientists, 42(1):3, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1986:SSR

[Bro86d] Harrison Brown. In search of a safe reactor. Bulletin of the Atomic Scientists, 42(8):3, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mac86b].

Brown:1986:MSS

[Bro86e] Harrison Brown. A myth spoils the summit. Bulletin of the Atomic Scientists, 42(10):3, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Win87, Fau87].

Brown:1986:TSPa

[Bro86f] Harrison Brown. Toward a strategy for peace — III. Bulletin of the Atomic Scientists, 42(2):3, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1986:TSPb

[Bro86g] Harrison Brown. Toward a strategy for peace — IV. Bulletin of the Atomic Scientists, 42(5):3, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1986:TSPc

[Bro86h] Harrison Brown. Toward a strategy for peace, V. Bulletin of the Atomic Scientists, 42(9):3, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1986:WHD

[Bro86i] Harrison Brown. What happened to diplomacy? Bulletin of the Atomic Scientists, 42(6):3, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1987:DLS

[Bro87a] Harrison Brown. Draw the line at Star Wars. Bulletin of the Atomic Scientists, 43(1):3, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1987:LSE

[Bro87b] Lowell Brown. Letter: Secretary of ethics? Bulletin of the Atomic Scientists, 43(2):63, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1987:PUC

[Bro87c] George E. Brown, Jr. Pentagon usurps civilian space program. Bulletin of the Atomic Scientists, 43(9):26–31, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brooks:1988:GTF

[Bro88a] Karen Brooks. Gorbachev tries the family farm. Bulletin of the Atomic Scientists, 44(10):26–29, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brown:1988:RTT

[Bro88b] Paul S. Brown. Round two of test ban flap: No I didn’t. Bulletin of the Atomic Scientists, 44(10):5–6, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eve88a, Eve88b].

Brociner:1989:LSB

[Bro89a] Ken Brociner. Letter: Solo book divisive. Bulletin of the Atomic Scientists, 45(4):53, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Swe89a].

Brown:1989:CDO

[Bro89b] George E. Brown, Jr. Courting disaster in orbit. Bulletin of the Atomic Scientists, 45(3):7–9, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brower:1990:BNN

[Bro90] Michael Brower. B-2: New numbers, old arguments. Bulletin of the Atomic Scientists, 46(5):25–29, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brooks:1993:LSE

[Bro93] Alexandra Brooks. Letter: Shut ’em down. Bulletin of the Atomic Scientists, 49(7):55, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pop93].

Brown:2005:JDB

[Bro05] Andrew Brown. J. D. Bernal: the Sage of Science. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2005. ISBN 0-19-851544-8. xiv + 562 + 16 pp. LCCN QC16.B457 B76

2005. URL http://site.ebrary.com/lib/stanford/Doc?id=
10177848;   http://www.loc.gov/catdir/toc/ecip0517/2005023505.
html;   http://www.myilibrary.com?id=75890.

Brougher:2013:ANC

[Bro13] Kerry Brougher. Art and nuclear culture. Bulletin of the Atomic Scientists, 69(6):11–18, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/6/11.

Bronson:2016:PSM

[Bro16] Rachel Bronson. Power shift in the Middle East. Bulletin of the Atomic Scientists, 72(1):10–15, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bronson:2020:BWB

[Bro20] Rachel Bronson. Buckle up: We are in for a bumpy ride. An interview with Royal Astronomer Martin Rees. Bulletin of the Atomic Scientists, 76(6):280–284, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1846406.

Bronson:2023:BSE

[Bro23] Rachel Bronson. Bulletin statement on the Energy Department’s Oppenheimer decision. Bulletin of the Atomic Scientists, 79(4):265–266, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2223086.

Brues:1947:ABC

[Bru47] Austin M. Brues. With the Atomic Bomb Casualty Commission in Japan. Bulletin of the Atomic Scientists, 3(6):143–144, 167, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brues:1948:MBR

[Bru48] Austin M. Brues. The “Mystery” of biological radiation effects. Bulletin of the Atomic Scientists, 4(11):341–342, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brues:1950:BBU

[Bru50] Austin M. Brues. Book: Constructive Uses of Atomic Energy, by S. C. Rothmann. Bulletin of the Atomic Scientists, 6(1):32, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bruneau:1952:SFE

[Bru52] Charles Bruneau. Some French experiences. Bulletin of the Atomic Scientists, 8(7):241–242, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Brues:1955:BRB

[Bru55a] Austin M. Brues. Book review: Hiroshima Diary, by Michihiko Hachiya. Bulletin of the Atomic Scientists, 11(7):267, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brues:1955:NER

[Bru55b] Austin M. Brues. The new emotionalism in research. Bulletin of the Atomic Scientists, 11(9):344–345, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bruyn:1955:PU

[Brü55c] Kathleen Brüyn. Power unlimited. Bulletin of the Atomic Scientists, 11(3):88–91, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brues:1957:BBE

[Bru57] Austin M. Brues. Books: The Effect of Exposure to the Atomic Bombs on Pregnancy Termination in Hiroshima and Nagasaki, by J. V. Neel and W. J. Schull. Bulletin of the Atomic Scientists, 13(10):372, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brues:1958:RMB

[Bru58] Austin M. Brues. Radiation and man: Basic data: Somatic effects of radiation. Bulletin of the Atomic Scientists, 14(1):12–14, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brues:1959:BBW

[Bru59] Austin M. Brues. Books: Nine Who Survived Hiroshima and Nagasaki, by Robert Trumbull. Bulletin of the Atomic Scientists, 15(4):173, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brunson:1972:CLG

[Bru72] G. S. Brunson. Communications: The language gap. Bulletin of the Atomic Scientists, 28(1):3–4, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Tam71a, Cha72].

Brush:1976:CSC

[Bru76] Stephen G. Brush. Can science come out of the laboratory now? Bulletin of the Atomic Scientists, 32(4):40–43, April 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ste76c, She77, Wal77].

Brumfield:1982:SUP

[Bru82] William C. Brumfield. The Soviet Union: Post-war architecture and planning. Bulletin of the Atomic Scientists, 38(3):14–17, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brunswick:1988:LGM

[Bru88] Les Brunswick. Letter: Good missile defense. Bulletin of the Atomic Scientists, 44(9):45, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rog88a].

Brugioni:1990:RDK

[Bru90] Dino A. Brugioni. Report: Declassified: The Kyshtym connection. Bulletin of the Atomic Scientists, 46(2):12, March 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bruce:1996:LNC

[Bru96] Frank R. Bruce. Letter: No Chernobyls here. Bulletin of the Atomic Scientists, 52(6):63, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano96r].

Brune:2015:OSM

[Bru15] William H. Brune. The ozone story: a model for addressing climate change? Bulletin of the Atomic Scientists, 71(1):75– 84, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brzezinski:1982:REW

[Brz82] Zbigniew Brzezinski. Rethinking East–West relations. Bulletin of the Atomic Scientists, 38(5):5–8, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brzoska:1987:PII

[Brz87] Michael Brzoska. Profiteering on the Iran–Iraq war. Bulletin of the Atomic Scientists, 43(5):42–45, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brzoska:1989:YS

[Brz89a] Michael Brzoska. At your service. Bulletin of the Atomic Scientists, 45(6):34, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brzoska:1989:BGE

[Brz89b] Michael Brzoska. Behind the German export scandals. Bulletin of the Atomic Scientists, 45(6):32–35, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Brzoska:2012:CCM

[Brz12] Michael Brzoska. Climate change and the military in China, Russia, the United Kingdom, and the United States. Bulletin of the Atomic Scientists, 68(2):43–54, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/68/2/43.

Bethe:1947:GSA

[BS47] H. A. Bethe and H. S. Sack. German scientists in Army employment: II. A protest. Bulletin of the Atomic Scientists, 3(2):65, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bardin:1957:RPA

[BS57] I. P. Bardin and M. M. Sisakyan. Resolution of the Presidium of the Academy of Sciences of the USSR. Bulletin of the Atomic Scientists, 13(9):316, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Best:1967:LEV

[BS67] Norman E. Best and Frederick Schwartz. Letter to the Editor: On Vietnam. Bulletin of the Atomic Scientists, 23(1):32–33, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab66e].

Bulkeley:1985:LSS

[BS85] Rip Bulkeley and Stephen Shenfield. Letter: Soviets and Star Wars. Bulletin of the Atomic Scientists, 41(9):53–54, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [She85b].

Block:1997:LMD

[BS97] Ellery Block and Stephen Schwartz. Letters: Missile defenses have a long history. Bulletin of the Atomic Scientists, 53(1):68– 69, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch96].

Burnie:2001:JNT

[BS01] Shaun Burnie and Aileen Mioko Smith. Japan’s nuclear twilight zone. Bulletin of the Atomic Scientists, 57(3):58–62, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/3/58.

Bird:2005:APT

[BS05] Kai Bird and Martin J. Sherwin. American Prometheus: the triumph and tragedy of J. Robert Oppenheimer. Alfred A. Knopf, New York, NY, USA, 2005. ISBN 0-375-72626-8, 0-375-41202-6. xiii + 721 + 32 pp. LCCN QC16.O62 B57 2005.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
;   http://site.ebrary.com/lib/ucsc/Doc?id=10078784.

Brock:2012:PPR

[BS12] Terry A. Brock and Sami S. Sherbini. Principles in practice: Radiation regulation and the NRC. Bulletin of the Atomic Scientists, 68(3):36–43, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/3/36.

Butler:2006:WJ

[BSS06] Kenley Butler, Sammy Salama, and Leonard S. Spector. Where is the justice? Bulletin of the Atomic Scientists, 62(6):25–34, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/6/25.

Barletta:2002:KTA

[BST02] Michael Barletta, Amy Sands, and Jonathan B. Tucker. Keeping track of anthrax: The case for a biosecurity convention. Bulletin of the Atomic Scientists, 58(3):57–62, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/3/57.

Beilenson:1982:CNN

[BT82] Laurence W. Beilenson and Kosta Tsipis. Commentary: New nuclear strategy. Bulletin of the Atomic Scientists, 38(7):60–61, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Car82b, Tsi82b].

Berrios:1994:PTO

[BT94] Rubén Berríos and Lillian Thomas. Perspective: Taking orders from Little Havana. Bulletin of the Atomic Scientists, 50(5):20– 21, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction to photograph credit [Ano94s].

Bunn:1995:LIE

[BT95] George Bunn and Roland M. Timerbaev. Letter: Indefinite extension, yes. Bulletin of the Atomic Scientists, 51(3):3, 68, May/ June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See opposing view [Toy95].

Buzzatti-Traverso:1966:LEL

[BTCSC^{+}66] A. Buzzatti-Traverso, L. Cavalli-Sforza, M. Cini, G. Cortini, G. Giacometti, F. Graziosi, and B. Vitale. Letter to the Editor: Letter from Italy on Vietnam. Bulletin of the Atomic Scientists, 22(1):38, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Buck:1950:FTA

[Buc50] J. Lossing Buck. Fact and theory about China’s land. Bulletin of the Atomic Scientists, 6(12):365–368, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Buchan:1960:FWD

[Buc60] Alastair Buchan. The future of Western deterrent power. III. A view from the United Kingdom. Bulletin of the Atomic Scientists, 16(7):277–282, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Buchsbaum:1977:LEB

[Buc77] Otto Buchsbaum. Letter: Endangered Brazil. Bulletin of the Atomic Scientists, 33(8):2, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Buchanan:1979:CPC

[Buc79] Todd Buchanan. Commentary: President Carter: Why the MX? Bulletin of the Atomic Scientists, 35(8):60–61, October 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Buchan:1981:AMM

[Buc81] Glenn C. Buchan. The anti-MAD mythology. Bulletin of the Atomic Scientists, 37(4):13–17, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gra81a, BCB81, Cou81, Chi82].

Buchheim:1982:CUS

[Buc82] Robert W. Buchheim. Commentary: U.S./USSR Consultative Commission. Bulletin of the Atomic Scientists, 38(3):49, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cla81].

Buchan:1983:VS

[Buc83] Glenn C. Buchan. The verification spectrum. Bulletin of the Atomic Scientists, 39(9):16–19, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Budjeryn:2022:DSD

[Bud22] Mariana Budjeryn. Distressing a system in distress: global nuclear order and Russia’s war against Ukraine. Bulletin of the Atomic Scientists, 78(6):339–346, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2022.
2132742.

Bugliarello:1978:TM

[Bug78] George Bugliarello. A technological magistrature. Bulletin of the Atomic Scientists, 34(1):34–37, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bugos:2022:DCU

[Bug22] Shannon Bugos. Despite challenges, US–Russian nuclear arms control has its benefits. Bulletin of the Atomic Scientists, 78 (6):334–338, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2132741.

Bukharin:1993:SLB

[Buk93] Oleg Bukharin. Soft landing for bomb uranium. Bulletin of the Atomic Scientists, 49(7):44–49, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bal93].

Bukharin:1997:DEF

[Buk97a] Oleg Bukharin. Disquiet on the Eastern Front. Bulletin of the Atomic Scientists, 53(3):41–46, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bukharin:1997:PPL

[Buk97b] Oleg Bukharin. A possible program …. Bulletin of the Atomic Scientists, 53(3):44–45, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bukharin:1997:TPW

[Buk97c] Oleg Bukharin. Tritium, pits, and warheads. Bulletin of the Atomic Scientists, 53(3):43, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bukharin:2002:MFL

[Buk02] Oleg Bukharin. Making fuel less tempting. Bulletin of the Atomic Scientists, 58(4):44–49, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/44.

Bukharin:2003:ACM

[Buk03] Oleg Bukharin. Article commentary: More transparency needed. Bulletin of the Atomic Scientists, 59(2):25–26, March 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/2/25.

Bull:1967:BBI

[Bul67] Hedley Bull. Books: Arms and Influence, by Thomas C. Schelling. Bulletin of the Atomic Scientists, 23(3):25–26, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bulkeley:1983:MET

[Bul83] Rip Bulkeley. Myths of Edward Teller. Bulletin of the Atomic Scientists, 39(8):64–65, October 1983. CODEN BASIAP. ISSN 0096-3402. See [vH83a].

Staff:2001:CYW

[Bul01a] Bulletin Staff. Consider yourself warned. Bulletin of the Atomic Scientists, 57(6):40–41, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/40.

Staff:2001:DWM

[Bul01b] Bulletin Staff. Defense when money is no object. Bulletin of the Atomic Scientists, 57(5):36–37, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/36.

Staff:2001:TVR

[Bul01c] Bulletin Staff. Threat v. reality. Bulletin of the Atomic Scientists, 57(4):36–37, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/4/36.

Staff:2002:YA

[Bul02a] Bulletin Staff. 60 years ago. … Bulletin of the Atomic Scientists, 58(6):54–55, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/54.

Staff:2002:BS

[Bul02b] Bulletin Staff. Battle stats. Bulletin of the Atomic Scientists, 58 (5):38–39, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/5/38.

Staff:2002:GIA

[Bul02c] Bulletin Staff. Going it alone. Bulletin of the Atomic Scientists, 58(4):36–37, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/4/36.

Staff:2002:HBC

[Bul02d] Bulletin Staff. The history of the Bulletin Clock. Bulletin of the Atomic Scientists, 58(2):36–37, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/36.

Staff:2002:SSS

[Bul02e] Bulletin Staff. It’s a Scud, Scud, Scud world. Bulletin of the Atomic Scientists, 58(1):36–37, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/36.

Staff:2002:PMD

[Bul02f] Bulletin Staff. Plutonium memorial design contest. Bulletin of the Atomic Scientists, 58(3):39, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/3/39.

Staff:2003:GRU

[Bul03a] Bulletin Staff. Guns R U.S. Bulletin of the Atomic Scientists, 59(1):38–39, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/59/1/38.

Staff:2003:NKL

[Bul03b] Bulletin Staff. North Korea: Less than meets the eye. Bulletin of the Atomic Scientists, 59(2):38–39, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/2/38.

Staff:2004:DC

[Bul04a] Bulletin Staff. Doomsdays for the 21st Century. Bulletin of the Atomic Scientists, 60(1):38–39, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/1/38.

Staff:2004:MTM

[Bul04b] Bulletin Staff. More than minutes: Announcing the new Bulletin contest! Bulletin of the Atomic Scientists, 60(5):36–37, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/5/36.

Staff:2005:BN

[Bul05] Bulletin Staff. Book notes. Bulletin of the Atomic Scientists, 61 (2):69, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/2/69.

Staff:2007:W

[Bul07a] Bulletin Staff. At the workbench. Bulletin of the Atomic Scientists, 63(4):46–47, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/4/46.

Staff:2007:IBT

[Bul07b] Bulletin Staff. Interview: Bruce Tarter. Bulletin of the Atomic Scientists, 63(4):24–29, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/4/24.

Staff:2007:PC

[Bul07c] Bulletin Staff. The People’s clock. Bulletin of the Atomic Scientists, 63(1):20–21, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/20.

Staff:2007:STI

[Bul07d] Bulletin Staff. The stealth threat: an interview with K. Eric Drexler. Bulletin of the Atomic Scientists, 63(1):55–58, January/ February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
1/55.

Staff:2008:APP

[Bul08a] Bulletin Staff. Aligning public and private interests to tackle climate change. Bulletin of the Atomic Scientists, 64(3):8–9, July/ August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/64/
3/8.

Staff:2008:BEN

[Bul08b] Bulletin Staff. Biodefense efforts need transparency. Bulletin of the Atomic Scientists, 64(5):6, November/December 2008. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/64/5/6.

Staff:2008:BBU

[Bul08c] Bulletin Staff. Building better U.S.–Russian relations. Bulletin of the Atomic Scientists, 64(1):7–8, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/1/7.

Staff:2008:BBa

[Bul08d] Bulletin Staff. The Bulletin board. Bulletin of the Atomic Scientists, 64(3):46–47, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/3/46.

Staff:2008:BBb

[Bul08e] Bulletin Staff. The Bulletin board. Bulletin of the Atomic Scientists, 64(4):58–59, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/4/58.

Staff:2008:BBc

[Bul08f] Bulletin Staff. The Bulletin board. Bulletin of the Atomic Scientists, 64(5):42–43, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/5/42.

Staff:2008:Ca

[Bul08g] Bulletin Staff. Contributors. Bulletin of the Atomic Scientists, 64(1):2–3, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/1/2.

Staff:2008:Cb

[Bul08h] Bulletin Staff. Contributors. Bulletin of the Atomic Scientists, 64(2):2–3, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/2/2.

Staff:2008:Cc

[Bul08i] Bulletin Staff. Contributors. Bulletin of the Atomic Scientists, 64(3):4–5, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/3/4.

Staff:2008:Cd

[Bul08j] Bulletin Staff. Contributors. Bulletin of the Atomic Scientists, 64(4):2–3, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/4/2.

Staff:2008:Ce

[Bul08k] Bulletin Staff. Contributors. Bulletin of the Atomic Scientists, 64(5):2–3, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/5/2.

Staff:2008:IJF

[Bul08l] Bulletin Staff. Interview: Jacqueline Fletcher. Bulletin of the Atomic Scientists, 64(3):10–13, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/3/10.

Staff:2008:IJR

[Bul08m] Bulletin Staff. Interview: John Rowe. Bulletin of the Atomic Scientists, 64(4):9–13, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/4/9.

Staff:2008:IPS

[Bul08n] Bulletin Staff. Interview: Peter Sellars. Bulletin of the Atomic Scientists, 64(1):10–15, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/1/10.

Staff:2008:ISR

[Bul08o] Bulletin Staff. Interview: Spencer R. Weart. Bulletin of the Atomic Scientists, 64(2):9–12, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/2/9.

Staff:2008:P

[Bul08p] Bulletin Staff. Passings. Bulletin of the Atomic Scientists, 64(5): 31–33, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/64/5/31.

Staff:2008:SNW

[Bul08q] Bulletin Staff. Sorting out nuclear waste. Bulletin of the Atomic Scientists, 64(4):7–8, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/4/7.

Staff:2008:TAN

[Bul08r] Bulletin Staff. Talking about nuclear energy. Bulletin of the Atomic Scientists, 64(2):7, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/2/7.

Staff:2008:TRP

[Bul08s] Bulletin Staff. Thomas R. Pickering. Bulletin of the Atomic Scientists, 64(5):7–11, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/5/7.

Staff:2009:BOV

[Bul09a] Bulletin Staff. Backing Obama’s vision for a nuclear-weapon-free world. Bulletin of the Atomic Scientists, 65(4):1–4, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/4/1.

Staff:2009:BISa

[Bul09b] Bulletin Staff. The Bulletin interview: Sergio Duarte. Bulletin of the Atomic Scientists, 65(3):4–12, May 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/65/3/4.

Staff:2009:BISb

[Bul09c] Bulletin Staff. The Bulletin interview: Stephen Schneider. Bulletin of the Atomic Scientists, 65(4):5–13, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/4/5.

Staff:2009:CNE

[Bul09d] Bulletin Staff. Confronting nuclear energy’s proliferation problem. Bulletin of the Atomic Scientists, 65(2):1–3, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/2/1.

Staff:2009:GUC

[Bul09e] Bulletin Staff. Getting U.S. climate policy right. Bulletin of the Atomic Scientists, 65(3):1–3, May 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/65/3/1.

Staff:2009:IME

[Bul09f] Bulletin Staff. Interview: Mohamed ElBaradei. Bulletin of the Atomic Scientists, 65(5):1–9, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/65/5/1.

Staff:2009:JH

[Bul09g] Bulletin Staff. Jennifer Hutchings. Bulletin of the Atomic Scientists, 65(1):4–12, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/65/1/4.

Staff:2009:PUN

[Bul09h] Bulletin Staff. A prescription for U.S. nuclear weapons policy. Bulletin of the Atomic Scientists, 65(1):1–3, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/1/1.

Staff:2009:RM

[Bul09i] Bulletin Staff. Richard Meserve. Bulletin of the Atomic Scientists, 65(2):4–12, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/65/2/4.

Staff:2010:IGC

[Bul10] Bulletin Staff. Interview: George Church. Bulletin of the Atomic Scientists, 66(1):34–40, January 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/66/1/34.

Bunche:1979:PNP

[Bun79] Ralph J. Bunche. Perspective: Nobel Peace Laureates: 1950. Bulletin of the Atomic Scientists, 35(1):1, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 11 December 1950, Oslo, Norway.

Bundy:1982:NFU

[Bun82] McGeorge Bundy. ‘No First Use’ needs careful study. Bulletin of the Atomic Scientists, 38(6):6–8, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bundy:1983:MBY

[Bun83] McGeorge Bundy. MX: not buried yet. Bulletin of the Atomic Scientists, 39(6):14–15, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bundy:1989:ACC

[Bun89] McGeorge Bundy. Arms control criteria: Negotiating skills needed. Bulletin of the Atomic Scientists, 45(4):33–34, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bunn:1998:GOA

[Bun98] Matthew Bunn. Guest opinion: Act now, Mr. President. Bulletin of the Atomic Scientists, 54(2):4, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burton:1947:RCB

[Bur47] Milton Burton. Radiation chemistry: A brief history and forecast. Bulletin of the Atomic Scientists, 3(12):366–369, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burgers:1948:LEV

[Bur48] J. M. Burgers. A letter from Europe [to Victor F. Weisskopf]. Bulletin of the Atomic Scientists, 4(1):6, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burgers:1950:LEC

[Bur50] J. M. Burgers. Letter to the Editor: Chicago atomic scientists offer free legal counsel to AEC employees. Bulletin of the Atomic Scientists, 6(3):79, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burtt:1957:VPS

[Bur57] E. A. Burtt. The value presuppositions of science. Bulletin of the Atomic Scientists, 13(3):99–106, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Man57].

Burkhardt:1962:DME

[Bur62a] Gerd Burkhardt. Disarmament in Middle Europe. Bulletin of the Atomic Scientists, 18(5):32–33, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burkhardt:1962:GSS

[Bur62b] Gerd Burkhardt. German scientists speak up. Bulletin of the Atomic Scientists, 18(6):45–46, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burhoe:1965:BBV

[Bur65] Ralph Wendell Burhoe. Books: My View of the World, by Erwin Schrödinger; The Relevance of Science, Creation and Cosmogony, by C. F. von Weizsäcker. Bulletin of the Atomic Scientists, 21 (7):33–35, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burkhardt:1966:SEA

[Bur66] G. Burkhardt. Science education in Africa. Bulletin of the Atomic Scientists, 22(2):46–48, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burstein:1972:CEM

[Bur72] Rabbi Samuel M. Burstein. Communications: Ecology and morality. Bulletin of the Atomic Scientists, 28(3):3–4, March 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Whi72].

Burhop:1974:SSA

[Bur74] E. H. S. Burhop. Scientists and soldiers: America’s ‘Jason Group’ looks back on its Vietnam involvement. Bulletin of the Atomic Scientists, 30(9):4–5, 7–8, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Dre74, Kuh75, Spe75].

Burns:1976:TAU

[Bur76] E. L. M. Burns. Thinking about the unthinkable. Bulletin of the Atomic Scientists, 32(1):42–43, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burcharth:1987:DBU

[Bur87] Martin Burcharth. Danes bristle at U.S. radar plans. Bulletin of the Atomic Scientists, 43(5):11–13, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burton:1988:LGA

[Bur88] John Burton. Letter: Gorbachev augments security. Bulletin of the Atomic Scientists, 44(1):54, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shu87].

Burck:1990:LDC

[Bur90a] Gordon Burck. Letter: Draft of chemical accord misread. Bulletin of the Atomic Scientists, 46(10):51, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bur90b].

Burck:1990:LLT

[Bur90b] Gordon Burck. Letter: Loophole threatens chemical weapons ban. Bulletin of the Atomic Scientists, 46(9):46, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap90b] and correction [Bur90a].

Burck:1990:BRC

[Bur90c] Gordon M. Burck. Book review: Chemical fog: Chemical Weaponry: A Continuing Challenge, by Edward M. Spiers. Bulletin of the Atomic Scientists, 46(8):43–44, October 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burnie:2005:DMB

[Bur05] Shaun Burnie. Detonating the mind bomb. Bulletin of the Atomic Scientists, 61(3):62–63, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/62.

Burroughs:2006:FH

[Bur06] William J. Burroughs. Feeling the heat. Bulletin of the Atomic Scientists, 62(6):53–54, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/6/53.

Burr:2009:BHU

[Bur09] William Burr. A brief history of U.S.–Iranian nuclear negotiations. Bulletin of the Atomic Scientists, 65(1):21–34, January/ February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/65/
1/21.

Week:1946:SDU

[Bus46] Business Week. Science dons a uniform. Bulletin of the Atomic Scientists, 2(9–10):11, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bush:1947:RRD

[Bus47] Vannevar Bush. Replies to Ridenour: Dangers to research, if recognized, can be avoided. Bulletin of the Atomic Scientists, 3 (8):228, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bushnell:1985:BRB

[Bus85] John Bushnell. Book reviews: Cold War, Cold Peace, by Bernard A. Weisberger; The USSR in Third World Conflicts, by Bruce D. Porter; Rethinking the Soviet Experience, by Stephen F. Cohen. Bulletin of the Atomic Scientists, 41(10):46–48, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bushev:1994:TKP

[Bus94] Aleksandr Bushev. Turkmenistan: A kind of prosperity. Bulletin of the Atomic Scientists, 50(1):44–47, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Butler:1973:CEC

[But73] Willis Butler. Communications: Ehrlich–Commoner debate. Bulletin of the Atomic Scientists, 29(1):2–3, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [EH72a, Com72].

Butler:1984:CNW

[But84] Brendan Butler. Commentary: Nuclear war on the farm. Bulletin of the Atomic Scientists, 40(10):33–34, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste84].

Butler:1997:TEA

[But97a] George Lee Butler. Time to end the age of nukes. Bulletin of the Atomic Scientists, 53(2):33–36, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Butler:1997:RWL

[But97b] Nicola Butler. Report: Weapons labs: Sharing secrets. Bulletin of the Atomic Scientists, 53(1):11–12, January/February 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Butler:1998:VR

[But98] Lee Butler. A voice of reason. Bulletin of the Atomic Scientists, 54(3):58–61, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kri98].

Butler:1999:CC

[But99] Judy Butler. Congressional cowardice. Bulletin of the Atomic Scientists, 55(3):74–75, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/3/74.

Butler:2000:ZT

[But00] Lee Butler. Zero tolerance. Bulletin of the Atomic Scientists, 56(1):20–21, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/56/1/20.

Butler:2004:BCA

[But04] Nicola Butler. Britain: Comrades in arms. Bulletin of the Atomic Scientists, 60(4):15–16, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/4/15.

Bidwai:1997:ACL

[BV97] Praful Bidwai and Achin Vanaik. After the CTB … India’s intentions. Bulletin of the Atomic Scientists, 53(2):49–50, March/ April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kar97].

Bidwai:1998:VPB

[BV98] Praful Bidwai and Achin Vanaik. A very political bomb. Bulletin of the Atomic Scientists, 54(4):50–52, July/August 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Blechman:2010:URZ

[BV10] Barry Blechman and Jonas Vaicikonis. Unblocking the road to zero: US–Russian cooperation on missile defenses. Bulletin of the Atomic Scientists, 66(6):25–35, November/December 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/6/25.

BenOuagrham-Gormley:2020:FMW

[BV20] Sonia Ben Ouagrham-Gormley and Kathleen M. Vogel. Follow the money: What the sources of Jiankui He’s funding reveal about what Beijing authorities knew about illegal CRISPR babies, and when they knew it. Bulletin of the Atomic Scientists, 76(4):192–199, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1780726.

Bode:2021:SCM

[BV21] Leticia Bode and Emily Vraga. The Swiss cheese model for mitigating online misinformation. Bulletin of the Atomic Scientists, 77(3):129–133, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1912170.

Beyea:1982:CRM

[BvH82] Jan Beyea and Frank von Hippel. Containment of a reactor meltdown. Bulletin of the Atomic Scientists, 38(7):52–59, August/ September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bellamy:1959:NSA

[BW59] Albert W. Bellamy and Stafford L. Warren. National survival in the Atomic Age. Bulletin of the Atomic Scientists, 15(9):390–392, 397, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ewi60].

Brown:1977:LLH

[BW77] Martin Brown and Alvin M. Weinberg. Letter: The linear hypothesis. Bulletin of the Atomic Scientists, 33(7):3, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei77b].

Barnea:1980:CNE

[BW80] Joseph Barnea and Alvin M. Weinberg. Commentary: Is nuclear energy necessary? Bulletin of the Atomic Scientists, 36(7):62– 63, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei80a].

Bedell:1983:CMD

[BW83] Douglas H. Bedell and Edwin J. Walsh. Commentary: On ‘Meltdown of democracy’. Bulletin of the Atomic Scientists, 39(7): 57–58, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wal83a].

Barkenbus:1984:BND

[BW84] Jack N. Barkenbus and Alvin M. Weinberg. Breaking the nuclear deadlock: Defense-protected build-down. Bulletin of the Atomic Scientists, 40(8):18–23, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lyn84].

Berkhout:1994:RNF

[BW94] Frans Berkhout and William Walker. Report: Nuclear fuel: Atlantic impasse. Bulletin of the Atomic Scientists, 50(5):15–17, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burger:2018:HFF

[BW18] Michael Burger and Jessica Wentz. Holding fossil fuel companies accountable for their contribution to climate change: Where does the law stand? Bulletin of the Atomic Scientists, 74 (6):397–403, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1533217.

Barkenbus:1985:SWS

[BWA85] Jack N. Barkenbus, Alvin M. Weinberg, and Marcelo Alonso. Storing the world’s spent nuclear fuel. Bulletin of the Atomic Scientists, 41(10):34–37, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bas86].

Bethe:1980:CSS

[BWH80] Hans A. Bethe, Richard Wilson, and Andrew Hull. Commentary: Scientists for SENSE. Bulletin of the Atomic Scientists, 36(7): 64, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bey80a].

Black:2021:UWU

[BWL21] Stephen Black, Henrietta Wilson, and Filippa Lentzos. UN-SCOM’s work to uncover Iraq’s illicit biological weapons program: A primer. Bulletin of the Atomic Scientists, 77(4):199– 202, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1941592.

Buchanan:1980:CDP

[BWS^{+}80] Todd Matthew Buchanan, Adrienne Elaine White, Bradley Penn Sturges, Thomas John Novak, Robert Bruce Marshall, and Ada Sheryl Cooper. Commentary: Dear President Carter and Members of Congress. Bulletin of the Atomic Scientists, 36(5): 65, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Byers:1987:LNT

[Bye87] Bruce A. Byers. Letter: Nuclear taxonomy. Bulletin of the Atomic Scientists, 43(1):57, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Byers:1988:LFC

[Bye88] Bruce A. Byers. Letter: Facts changes opinions. Bulletin of the Atomic Scientists, 44(9):46, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [CN88].

Byers:1991:LES

[Bye91] Bruce A. Byers. Letter: Environmental security network. Bulletin of the Atomic Scientists, 47(5):45, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gle91, Deu91].

AlNahyan:2008:WGN

[bZAAFC08] Abdullah bin Zayed Al Nahyan, Kamal J. Araj, Nabil Fahmy, and Somporn Chongkum. Why go nuclear? Bulletin of the Atomic Scientists, 64(4):14–19, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/4/14.

Cochran:1987:NWDa

[C^{+}87a] Thomas B. Cochran et al. Nuclear weapons databook. Volume 2. US nuclear warhead production. Nuclear weapons databook. Ballinger Publishing Company, Cambridge, MA, USA, 1987. ISBN 0-88730-124-X, 0-88730-125-8 (paperback). xiv + 223 pp. LCCN U264 .N827 1984 vol. 2.

Cochran:1987:NWDb

[C^{+}87b] Thomas B. Cochran et al. Nuclear weapons databook. Volume 3. US nuclear warhead facility profiles. Nuclear weapons databook. Ballinger Publishing Company, Cambridge, MA, USA, 1987. ISBN 0-88730-126-6, 0-88730-146-0 (paperback). xv + 132 pp. LCCN U264 .N827 1984 vol. 3.

Cochran:1989:NWD

[C^{+}89] Thomas B. Cochran et al. Nuclear weapons databook. Volume 4. Soviet nuclear weapons. Nuclear weapons databook. Harper and Row, Ballinger Division, New York, NY, USA, 1989. ISBN 0-88730-048-0, 0-88730-049-9 (paperback). xviii + 433 pp. LCCN U264 .N827 1984 vol. 4.

Charles:1984:EN

[CA84] Daniel Charles and David Albright. Europeanization of NATO. Bulletin of the Atomic Scientists, 40(9):45–46, November 1984.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cain:1987:RMB

[CA87] Stephen Alexis Cain and Gordon Adams. Reagan’s 1988 military budget. Bulletin of the Atomic Scientists, 43(2):50–52, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Caldicott:1998:LDD

[CA98a] Helen Caldicott and William M. Arkin. Letter: Dangerous DU. Bulletin of the Atomic Scientists, 54(1):68–69, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark97e].

Cortright:1998:SME

[CA98b] David Cortright and Samina Ahmed. Sanctions: Modify ’em. Bulletin of the Atomic Scientists, 54(5):22–24, September/ October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cabasso:1997:LTT

[Cab97] Jacqueline Cabasso. Letter: Trouble in Tahiti. Bulletin of the Atomic Scientists, 53(3):60, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Caesar:1973:CMI

[Cae73] Irving Caesar. Communications: Military–industrial complex. Bulletin of the Atomic Scientists, 29(2):2–3, February 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tsi72b].

Caesar:1974:LKD

[Cae74] Irving Caesar. Letter: Keep digging. Bulletin of the Atomic Scientists, 30(10):2, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [JL74].

Cahn:1971:SVP

[Cah71] Anne H. Cahn. A student’s view of Pugwash. Bulletin of the Atomic Scientists, 27(4):36–37, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cahn:1975:AWS

[Cah75] Anne Hessing Cahn. Have arms, will sell: Quantity is up and so is quality. Bulletin of the Atomic Scientists, 31(4):10–12, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cahn:1976:BSP

[Cah76] Anne Hessing Cahn. Books: Scientists, politicians and the bomb: A World Destroyed: The Atomic Bomb and the Grand Alliance, by Martin J. Sherwin. Bulletin of the Atomic Scientists, 32(4): 44–46, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cochran:1984:NWD

[CAH84] Thomas B. Cochran, William M. Arkin, and Milton M. Hoening. Nuclear weapons databook. Volume 1. U.S. nuclear forces and capabilities. Ballinger Publishing Company, Cambridge, MA, USA, 1984. ISBN 0-88410-172-X, 0-88410-173-8 (paperback). ???? pp. LCCN U264 .C6 1984.

Cahn:1994:BRN

[Cah94] Anne H. Cahn. Book review: Nitze’s list: Tension between Opposites: Reflections on the Practice and Theory of Politics, by Paul H. Nitze. Bulletin of the Atomic Scientists, 50(4):56–57, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cain:1988:CBA

[Cai88] Stephen Alexis Cain. Carlucci’s budget avoids hard choices. Bulletin of the Atomic Scientists, 44(4):44–46, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cain:1989:WRM

[Cai89] Stephen Alexis Cain. Washington report: Military budget: One more for the Gipper. Bulletin of the Atomic Scientists, 45(6):5–6, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cain:1990:RMB

[Cai90] Stephen Alexis Cain. Report: Military budget: Just a trim, please. Bulletin of the Atomic Scientists, 46(3):7–8, 51–52, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calingaert:1957:LEG

[Cal57] George Calingaert. Letter to the Editor: Galileo rescinded? Bulletin of the Atomic Scientists, 13(5):184, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dub56a] and response [Dub57a].

Calder:1960:LLH

[Cal60a] Ritchie Calder. Letter from London: As the hare said to the tortoise? Bulletin of the Atomic Scientists, 16(9):352–354, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1960:LLN

[Cal60b] Ritchie Calder. Letter from London: The non-nuclear club. Bulletin of the Atomic Scientists, 16(4):123–126, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1961:DIY

[Cal61a] Ritchie Calder. A do-it-yourself conference: Report from Rome. Bulletin of the Atomic Scientists, 17(10):425–426, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1961:LIN

[Cal61b] Ritchie Calder. Letter from Israel: New magic for old. Bulletin of the Atomic Scientists, 17(2):55–56, 77, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1962:BBB

[Cal62] Ritchie Calder. Books: The Birth of the Bomb, by R. W. Clark. Bulletin of the Atomic Scientists, 18(1):39–40, January 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1963:NRM

[Cal63] Ritchie Calder. News & reports: More food for Southeast Asia. Bulletin of the Atomic Scientists, 19(3):31–32, March 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1964:TI

[Cal64] Nigel Calder. Technology and the investor. Bulletin of the Atomic Scientists, 20(4):47–48, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1965:SC

[Cal65] Ritchie Calder. The speed of change. Bulletin of the Atomic Scientists, 21(10):2–5, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calder:1966:SCB

[Cal66] Nigel Calder. Supersonic confidence — a British view. Bulletin of the Atomic Scientists, 22(10):9, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calogero:1973:SES

[Cal73] Francesco Calogero. A scenario for effectiveness SALT negotiations. Bulletin of the Atomic Scientists, 29(6):16–22, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Caldwell:1980:CES

[Cal80] Dan Caldwell. CTB — an effective SALT substitute. Bulletin of the Atomic Scientists, 36(10):30–33, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Callahan:1982:BRB

[Cal82] Daniel Callahan. Book review: AAAS Professional Ethics Project, by Rosemary Chalk, Mark S. Frankel, and Sallie B. Chafer. Bulletin of the Atomic Scientists, 38(1):43, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Calogero:1984:BRB

[Cal84a] Francesco Calogero. Book reviews: Nuclear Weapons Databook: Volume 1: U.S. Nuclear Forces and Capabilities, by Thomas B. Cochran, William M. Arkin and Milton M. Hoenig; Arsenal: Understanding Weapons in the Nuclear Age, by Kosta Tsipis. Bulletin of the Atomic Scientists, 40(7):42–43, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano84g].

Calogero:1984:WME

[Cal84b] Francesco Calogero. Why missiles in Europe? Bulletin of the Atomic Scientists, 40(5):12–13, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Caldicott:1986:LHC

[Cal86a] Helen Caldicott. Letter: Helen Caldicott on tactics. Bulletin of the Atomic Scientists, 42(5):45–46, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Boy86a, SV86, Min86].

Caldicott:1986:PH

[Cal86b] Helen Caldicott. PSR history. Bulletin of the Atomic Scientists, 42(3):57, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano85m].

Calogero:1997:PFT

[Cal97] Francesco Calogero. Perspective: Fast-track the uranium deal. Bulletin of the Atomic Scientists, 53(6):20–21, November/ December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [YC98].

Calhoun:2017:DAP

[Cal17] Laurie Calhoun. Death from above: The perils of lethal drone strikes. Bulletin of the Atomic Scientists, 73(2):138–142, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1288459.

Calvelli:2024:PCV

[Cal24] Aidan G. Calvelli. The path to compulsory voting. Bulletin of the Atomic Scientists, 80(5):291–296, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2388464.

Cammer:1953:LEY

[Cam53] P. Cammer. Letter to the Editor: Your disgusting September issue. Bulletin of the Atomic Scientists, 9(10):393, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Campbell:1954:BRB

[Cam54] Kathleen Campbell. Book review: Atoms in the Family, by Laura Fermi, Chicago: University of Chicago Press, 1954. Bulletin of the Atomic Scientists, 10(10):393–394, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cammer:1960:LEM

[Cam60] P. Cammer. Letter to the Editor: Moving back the clock — pro. Bulletin of the Atomic Scientists, 16(3):80, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Camp:1976:LE

[Cam76] David B. Camp. Letter: Evolution. Bulletin of the Atomic Scientists, 32(7):3, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cambel:1977:LMC

[Cam77] Ali B. Cambel. Letter: ‘misleading’ critique. Bulletin of the Atomic Scientists, 33(1):2–3, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cas76].

Camilleri:1979:CAN

[Cam79] Joseph Camilleri. Commentary from Australia: Nuclear controversy in Australia: the uranium campaign. Bulletin of the Atomic Scientists, 35(4):40–44, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Campbell:1988:LLA

[Cam88] Robert L. Campbell. Letter: Lessons of atomic veterans. Bulletin of the Atomic Scientists, 44(7):59–60, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [La 88b].

Campbell:1989:BRB

[Cam89] Robert W. Campbell. Book review: The Economic Challenge of Perestroika, by Abel Aganbegyan. Bulletin of the Atomic Scientists, 45(4):47–48, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Campbell:1990:LUN

[Cam90] R. M. Campbell. Letter: U.S. needs more nuclear power, now. Bulletin of the Atomic Scientists, 46(5):44–45, June 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WF90].

Campbell:1995:LCW

[Cam95a] Robert Campbell. Letter: Credit where it’s due. Bulletin of the Atomic Scientists, 51(2):71, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Campbell:1995:LTA

[Cam95b] Robert L. Campbell. Letter: Teach all the madness. Bulletin of the Atomic Scientists, 51(4):76, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber95, Gol95c].

Causevic:2020:CNE

[CAM20] Amar Causevic and Ibrahim Al-Marashi. Can NATO evolve into a climate alliance treaty organization in the Middle East? Bulletin of the Atomic Scientists, 76(2):97–101, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1728981.

Cannan:1963:ABC

[Can63] R. Keith Cannan. Atomic Bomb Casualty Commission: The first fifteen years. Bulletin of the Atomic Scientists, 19(8):45– 48, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cannon:1970:BBF

[Can70] William B. Cannon. Books: Science and the Federal Patron, by Michael D. Reagan. Bulletin of the Atomic Scientists, 26(5): 45–47, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cantor:1972:CLD

[Can72] Arnold J. Cantor. Communications: On living dangerously. Bulletin of the Atomic Scientists, 28(4):3, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab72g].

CUSPG:1974:NNE

[Can74] Canadian and U. S. Pugwash Groups. Needed: A new ethic. Bulletin of the Atomic Scientists, 30(7):6–7, September 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Nat74].

CAPG:1979:PNS

[Can79] Canadian and American Pugwash Groups. From Pugwash, Nova Scotia: On nuclear disarmament. Bulletin of the Atomic Scientists, 35(7):7, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cochran:1988:MWU

[CAN88] Thomas B. Cochran, William M. Arkin, and Robert S. Norris. Making warheads: U.S. nuclear weapons production: an overview. Bulletin of the Atomic Scientists, 44(1):13–16, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://docs.nrdc.org/nuclear/
nuc_88010001a_79.pdf. See comment [Shi88].

Caputo:1977:SPP

[Cap77] Richard S. Caputo. Solar power plants: dark horse in the energy stable. Bulletin of the Atomic Scientists, 33(5):46–48, 50–56, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Capaccio:1998:BR

[Cap98] Tony Capaccio. Book review: The bombings reexamined: Prompt and Utter Destruction: President Truman and the Use of Atomic Bombs Against Japan, by J. Samuel Walker. Bulletin of the Atomic Scientists, 54(2):66–68, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Capozza:2002:DLD

[Cap02] K. L. Capozza. The Dew Line: Ditched drums and all. Bulletin of the Atomic Scientists, 58(1):14–16, January/February 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/1/14.

Carpenter:1949:CWR

[Car49a] Francis W. Carpenter. Conservation of world resources: A report on U.N. scientific conference. Bulletin of the Atomic Scientists, 5(11):313–314, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1949:UNAa

[Car49b] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 5(5):157, 160, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1949:UNAb

[Car49c] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 5(8–9):255–260, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1949:UNAc

[Car49d] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 5(11):320–322, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1949:UNAd

[Car49e] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 5(12):345–347, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1950:UNAa

[Car50a] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 6(1):18–20, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1950:UNAb

[Car50b] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 6(2):50, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1950:UNAc

[Car50c] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 6(3):90–91, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1950:UNAd

[Car50d] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 6(7):206, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1950:UNAe

[Car50e] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 6(11):342–343, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:1950:UNAf

[Car50f] Francis W. Carpenter. United Nations atomic energy news. Bulletin of the Atomic Scientists, 6(12):369–370, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carlson:1953:KF

[Car53] Anton J. Carlson. “Know the facts”. Bulletin of the Atomic Scientists, 9(9):331, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carter:1955:GPI

[Car55] T. C. Carter. The genetic problem of irradiated human populations. Bulletin of the Atomic Scientists, 11(10):362–363, 366, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carlson:1968:HJM

[Car68] E. A. Carlson. [H. J. Muller’s life]. Bulletin of the Atomic Scientists, 24(3):6, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carney:1969:ED

[Car69] David Carney. Economic development. Bulletin of the Atomic Scientists, 25(9):11–15, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carl:1972:LHS

[Car72] Ann Carl. The Lloyd Harbor Study Group intervention — a response. Bulletin of the Atomic Scientists, 28(6):31–36, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sai72].

Carter:1976:TST

[Car76] Jimmy Carter. Three steps toward nuclear responsibility. Bulletin of the Atomic Scientists, 32(8):8–14, October 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zim77].

Carroll:1977:LSS

[Car77] Anne Carroll. Letter: Sink or swim. Bulletin of the Atomic Scientists, 33(5):2, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rom77].

Carnesale:1981:CJB

[Car81] Albert Carnesale. Commentary: June 7 in Baghdad. Bulletin of the Atomic Scientists, 37(7):11–13, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carey:1982:GNP

[Car82a] Michael J. Carey. Growing up nuclear: Psychological fallout. Bulletin of the Atomic Scientists, 38(1):20–24, January 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carver:1982:NN

[Car82b] Field Marshall Lord Michael Carver. Nuclear nonsense. Bulletin of the Atomic Scientists, 38(4):5–7, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BT82].

Carey:1983:PSS

[Car83a] William D. Carey. Perspective: The secrecy syndrome. Bulletin of the Atomic Scientists, 39(7):2, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carlin:1983:BND

[Car83b] Jerome E. Carlin. Breaking the nuclear deadlock: Toward minimum deterrence. Bulletin of the Atomic Scientists, 39(10):50–52, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carver:1983:NFU

[Car83c] Michael Carver. No first use: a view from Europe. Bulletin of the Atomic Scientists, 39(3):22–26, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [TC83].

Carroll:1984:AWD

[Car84a] John E. Carroll. Agenda 1984: Water dampens U.S.–Canadian relations. Bulletin of the Atomic Scientists, 40(1):20–25, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho84a].

Carver:1984:WO

[Car84b] Field Marshal Lord Michael Carver. A window of opportunity. Bulletin of the Atomic Scientists, 40(5):10–12, May 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carroll:1989:ACS

[Car89] John E. Carroll. The acid challenge to security. Bulletin of the Atomic Scientists, 45(8):32–34, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [RC90].

Carter:1994:LUI

[Car94a] Luther J. Carter. Let’s use it. Bulletin of the Atomic Scientists, 50(3):42–44, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [LC94].

Carter:1994:RNP

[Car94b] Luther J. Carter. Report: Nuclear power: Selin on safety. Bulletin of the Atomic Scientists, 50(6):13–14, 57, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carter:1994:RNW

[Car94c] Luther J. Carter. Report: Nuclear waste: The Mescalero option. Bulletin of the Atomic Scientists, 50(5):11–13, September/ October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carter:1994:SVS

[Car94d] Luther J. Carter. Safety violations sampler. Bulletin of the Atomic Scientists, 50(6):14, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carter:1997:PTL

[Car97] Luther J. Carter. Perspective: It’s time to lay this waste to rest. Bulletin of the Atomic Scientists, 53(1):13–15, January/ February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carlson:2003:HUS

[Car03] Bengt Carlson. How Ulam set the stage: history has not given enough credit to the main man behind the H-bomb. Bulletin of the Atomic Scientists, 59(4):46–51, July/August 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/4/46;   http://
www.highbeam.com/doc/1G1-105163616.html.

Carr:2013:MAW

[Car13] Jeffrey Carr. The misunderstood acronym: Why cyber weapons aren’t WMD. Bulletin of the Atomic Scientists, 69(5):32– 37, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/5/32.

Carranza:2017:MNR

[Car17] Mario E. Carranza. Managing nuclear risk in South Asia. Bulletin of the Atomic Scientists, 73(1):64–66, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2016.1264219.

Cassel:1974:LGC

[Cas74] Hans M. Cassel. Letters: Genocide convention. Bulletin of the Atomic Scientists, 30(6):4, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1976:EST

[Cas76] Barry M. Casper. The experts’ silent trumpet. Bulletin of the Atomic Scientists, 32(7):24–25, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cam77].

Casper:1977:CCT

[Cas77a] Barry M. Casper. Congress and the cozy triangles: the case of energy. Bulletin of the Atomic Scientists, 33(5):5, May 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1977:LEN

[Cas77b] Barry M. Casper. Laser enrichment: a new path to proliferation? Bulletin of the Atomic Scientists, 33(1):28–41, January 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [GC77, Dic77, Cas77c].

Casper:1977:RTM

[Cas77c] Barry M. Casper. Response: Time for a moratorium. Bulletin of the Atomic Scientists, 33(6):56, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1977:SH

[Cas77d] Barry M. Casper. Scientists on Hill. Bulletin of the Atomic Scientists, 33(9):8–15, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1977:SAA

[Cas77e] Barry M. Casper. [Sidebar:] atomic approach to laser isotope separation. Bulletin of the Atomic Scientists, 33(1):32–, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1977:STM

[Cas77f] Barry M. Casper. [Sidebar:] ‘Tell me, Mr. Scientist … ’. Bulletin of the Atomic Scientists, 33(9):12, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cassidy:1977:BCE

[Cas77g] Harold G. Cassidy. Boundary conditions in energy and ecology. Bulletin of the Atomic Scientists, 33(3):31–32, March 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1978:NDO

[Cas78a] Barry M. Casper. New directions for OTA: ‘the more controversial, the better … ’. Bulletin of the Atomic Scientists, 34(2): 26–27, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1978:RRC

[Cas78b] Barry M. Casper. The rhetoric and reality of Congressional technology assessment. Bulletin of the Atomic Scientists, 34(2):20–31, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Casper:1984:AP

[Cas84] Barry M. Casper. An appeal to physicists. Bulletin of the Atomic Scientists, 40(8):9–13, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Caster:1985:LPS

[Cas85a] John H. Caster. Letter: Pressuring the Soviet Union. Bulletin of the Atomic Scientists, 41(9):54, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bla85a].

Castle:1985:LC

[Cas85b] Dennis H. Castle. Letter: Congratulations. Bulletin of the Atomic Scientists, 41(2):54, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AuC84, Ark84f].

Cassidy:1992:ULS

[Cas92] David C. Cassidy. Uncertainty: the life and science of Werner Heisenberg. W. H. Freeman, New York, NY, USA, 1992. ISBN 0-7167-2243-7. xii + 669 + 16 pp. LCCN QC16.H35 C37 1992; QC 16 .H35C37 1992. US$29.95.

Casper:2002:CWI

[Cas02] Monica J. Casper. Chemical weapons: Incineration Island. Bulletin of the Atomic Scientists, 58(2):17–19, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/2/17.

Caspary:2004:BAA

[Cas04] William R. Caspary. Not so bad after all? Bulletin of the Atomic Scientists, 60(2):83–84, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/2/83.

Cascio:2019:AEW

[Cas19] Jamais Cascio. The apocalypse: It’s not the end of the world. Bulletin of the Atomic Scientists, 75(6):269–272, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1680047.

Catherwood:1987:LNO

[Cat87] Hugh R. Catherwood. Letter: Number one in soybeans, missiles. Bulletin of the Atomic Scientists, 43(8):53, October 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tsi87].

Catherwood:1989:LNN

[Cat89] Hugh R. Catherwood. Letter: Numbers nonsense. Bulletin of the Atomic Scientists, 45(2):47, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RR88].

Cavers:1947:APV

[Cav47] David F. Cavers. Atomic power versus world security. Bulletin of the Atomic Scientists, 3(10):283–288, 302, October 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1948:NLU

[Cav48] David F. Cavers. New life for the UNAEC. Bulletin of the Atomic Scientists, 4(12):355–358, 362, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1950:IPI

[Cav50] David F. Cavers. An interim plan for international control of atomic energy. Bulletin of the Atomic Scientists, 6(1):13–16, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1952:ACD

[Cav52] David F. Cavers. Atomic controls in disarmament planning. Bulletin of the Atomic Scientists, 8(3):84–87, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1953:LMM

[Cav53] David F. Cavers. Legal measures to mitigate the economic impact of atomic attack. Bulletin of the Atomic Scientists, 9(7):269–272, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1955:ASE

[Cav55] David F. Cavers. The arms stalemate ends. Bulletin of the Atomic Scientists, 11(1):9–12, 18, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1956:ACU

[Cav56] David F. Cavers. Arms control in the United Nations: A decade of disagreement. Bulletin of the Atomic Scientists, 12(4):105–111, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1957:BBW

[Cav57] David Cavers. Books: Atomic Weapons and East–West Relations, by P. M. S. Blackett. Bulletin of the Atomic Scientists, 13(3):108– 109, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1958:BBD

[Cav58] David Cavers. Books: Inspection for Disarmament, edited by Seymour Melman. Bulletin of the Atomic Scientists, 14(9):394– 395, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1961:BBN

[Cav61] David F. Cavers. Books: Postwar Negotiations for Arms Control, by Bernhard G. Bechhoefer. Bulletin of the Atomic Scientists, 17 (10):436, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavers:1979:CEP

[Cav79] David F. Cavers. That Carter evacuation plan. Bulletin of the Atomic Scientists, 35(4):15–19, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavalieri:1982:TPN

[Cav82] Liebe F. Cavalieri. Twin perils: nuclear science and genetic engineering. Bulletin of the Atomic Scientists, 38(10):72–75, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [GC83].

Cavender:1990:BPB

[Cav90] Jane Cavender. Bulletin: Polish brain drain. Bulletin of the Atomic Scientists, 46(8):5, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cave:1992:RRN

[Cav92] Jane Cave. Report: Russia: No soft landing for academy. Bulletin of the Atomic Scientists, 48(2):8–9, 48, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cavallo:2004:OIP

[Cav04] Alfred Cavallo. Oil: The illusion of plenty. Bulletin of the Atomic Scientists, 60(1):??, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/1/20.

Cavallo:2005:OCE

[Cav05] Alfred J. Cavallo. Oil: Caveat empty. Bulletin of the Atomic Scientists, 61(3):16–18, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/16.

Cavallo:2006:PP

[Cav06] Alfred J. Cavallo. Playing the pump. Bulletin of the Atomic Scientists, 62(4):20–22, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/4/20.

Cavallo:2007:WT

[Cav07] Alfred J. Cavallo. Wishful thinking. Bulletin of the Atomic Scientists, 63(3):65–67, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/3/65.

Cavallo:2013:ERH

[Cav13] Alfred Cavallo. Elephant in the room: How OPEC sets oil prices and limits carbon emissions. Bulletin of the Atomic Scientists, 69(4):18–29, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/4/18.

Cohen:1979:CFN

[CB79] Bernard L. Cohen and Lawrence Badash. Commentary: A few numbers. Bulletin of the Atomic Scientists, 35(1):63–64, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bad78].

Cabasso:1995:WSL

[CB95] Jacqueline Cabasso and John Burroughs. The Western States Legal Foundation replies to Hugh Gusterson: End run around the NPT. Bulletin of the Atomic Scientists, 51(5):27–29, September/ October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gus95, Kat95].

Cohen:2006:ICT

[CB06] Avner Cohen and William Burr. Israel crosses the threshold. Bulletin of the Atomic Scientists, 62(3):22–30, May/June 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/3/22.

Cook:1995:LYM

[CBSF95] Gregory N. Cook, Jeremy M. Boak, and Kristin Shrader-Frechette. Letters: Yucca Mountain logic. Bulletin of the Atomic Scientists, 51(2):74–77, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SF94].

Choucri:2013:WCC

[CC13] Nazli Choucri and David D. Clark. Who controls cyberspace? Bulletin of the Atomic Scientists, 69(5):21–31, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/5/21.

Cori:1952:NNP

[CCD^{+}52] Carl F. Cori, Gerti T. Cori, C. J. Davisson, Joseph Erlanger, Philip S. Hench, H. J. Muller, William P. Murphy, Harold C. Urey, and George H. Whipple. Nine Nobel Prize-winners address Joliot-Curie. Bulletin of the Atomic Scientists, 8(5):131, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chilton:1972:MLT

[CCH72] Frank Chilton, Howard T. Coffey, and L. O. Hoppie. Magnetic levitation: Tomorrow’s transportation. Bulletin of the Atomic Scientists, 28(3):39–41, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Compton:1946:SCS

[CCMU46] A. H. Compton, Karl T. Compton, F. R. Moulton, and H. C. Urey. Scientists comment on State Dep’t report: Comments by scientists. Bulletin of the Atomic Scientists, 1(9):12, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Compton:1948:PSC

[CD48] Arthur H. Compton and Farrington Daniels. A poll of scientists at Chicago, July 1945. Bulletin of the Atomic Scientists, 4(2):44, 63, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1977:LAN

[CD77] Bernard L. Cohen and Samuel H. Day, Jr. Letter: Abundance of nuclear fuel. Bulletin of the Atomic Scientists, 33(4):2, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Colglazier:1980:UDN

[CD80] William Colglazier and Paul Doty. U.S. debates a new agency. Bulletin of the Atomic Scientists, 36(5):43–47, May 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Conway:1988:LWS

[CE88] Paul G. Conway and Robert D. English. Letter: Who said it? Bulletin of the Atomic Scientists, 44(8):54, October 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eng88].

Cohen:1985:LHN

[CEE85] Bernhard Cohen, Paul Ehrlich, and Anne Ehrlich. Letter: Hazards of nuclear power. Bulletin of the Atomic Scientists, 41(8):60– 61, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [EE85] and comment [Bey85].

Chayes:1977:SSA

[CET77] Abram Chayes, William Epstein, and Theodore B. Taylor. A surveillance satellite for all. Bulletin of the Atomic Scientists, 33 (1):7, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cousins:1946:BSR

[CF46] Norman Cousins and Thomas K. Finletter. A beginning for sanity: A review of the Acheson–Lilienthal report. Bulletin of the Atomic Scientists, 2(1–2):11–14, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carrillo-Flores:1974:WPC

[CF74] Antonio Carrillo-Flores. World Population Conference/1974: Toward a humane balance between life and death. Bulletin of the Atomic Scientists, 30(6):10–12, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1987:INA

[CF87] Avner Cohen and Benjamin Frankel. Israel’s nuclear ambiguity. Bulletin of the Atomic Scientists, 43(2):15–19, March 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [FP89].

Chamberlain:1981:CCP

[CFG^{+}81] Owen Chamberlain, Herman Feshbach, Donald Glaser, Sheldon Glashow, Leon Lederman, Francis Low, Philip Morrison, Edward Purcell, Victor Weisskopf, Jerome Wiesner, and Robert R. Wilson. Commentary: Concerned physicists. Bulletin of the Atomic Scientists, 37(8):65, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Open letter to US President Ronald Reagan.

Cochran:2010:TGB

[CFM^{+}10] Thomas B. Cochran, Harold A. Feiveson, Zia Mian, M. V. Ramana, Mycle Schneider, and Frank N. von Hippel. It’s time to give up on breeder reactors. Bulletin of the Atomic Scientists, 66(3):50–56, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/3/50.

Compton:1947:SCL

[CFU47] Karl T. Compton, Enrico Fermi, and Harold Urey. Scientists comment on Lilienthal. Bulletin of the Atomic Scientists, 3(3): 75, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carpenter:2004:HDB

[CG04a] Tom Carpenter and Clare Gilbert. Hanford: Don’t breathe the air. Bulletin of the Atomic Scientists, 60(3):24–30, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/3/24.

Cohen:2004:NNM

[CG04b] Avner Cohen and Thomas Graham, Jr. An NPT for nonmembers. Bulletin of the Atomic Scientists, 60(3):40–44, May/ June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/60/3/40.

Choucri:2012:LCH

[CG12] Nazli Choucri and Daniel Goldsmith. Lost in cyberspace: Harnessing the Internet, international relations, and global security. Bulletin of the Atomic Scientists, 68(2):70–77, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/2/70.

Chankin-Gould:2005:DES

[CGO05] Sarah Chankin-Gould and Ivan Oelrich. Double-edged shield. Bulletin of the Atomic Scientists, 61(3):36–41, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/3/36.

Cole:1954:PAE

[CH54a] Sterling Cole and Bourke Hickenlooper. Proposed Atomic Energy Act amendments. Bulletin of the Atomic Scientists, 10(6):227, 240, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cole:1954:SOM

[CH54b] Sterling Cole and Bourke B. Hickenlooper. Statement [on the Oppenheimer matter]. Bulletin of the Atomic Scientists, 10(5): 179, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffman:1971:CEE

[CH71] Richard B. Coffman and Garrett Hardin. Communications: Economics, ecology and political decisions. Bulletin of the Atomic Scientists, 27(4):2–4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Har70].

Chalonge:1952:SFE

[Cha52] Daniel Chalonge. Some French experiences. Bulletin of the Atomic Scientists, 8(7):238–240, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Chapman:1953:SPP

[Cha53] Dwight W. Chapman. Some psychological problems in civil defense. Bulletin of the Atomic Scientists, 9(7):280–282, 284, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chamberlain:1958:RMP

[Cha58a] Richard H. Chamberlain. Radiation and man: Peacetime radiation: Medical use of radiation: Risks and rewards. Bulletin of the Atomic Scientists, 14(1):39–40, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charpie:1958:BBN

[Cha58b] Robert A. Charpie. Books: On Nuclear Energy, by Donald J. Hughes. Bulletin of the Atomic Scientists, 14(8):336, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chapman:1960:ICI

[Cha60] Sydney Chapman. International cooperation and the IGY. Bulletin of the Atomic Scientists, 16(5):173–178, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chamberlain:1962:BBH

[Cha62] Owen Chamberlain. Books: The Hard Way to Peace, by Amitai Etzioni. Bulletin of the Atomic Scientists, 18(8):31, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chapman:1963:DWW

[Cha63a] Seville Chapman. Do we want fallout shelters? Bulletin of the Atomic Scientists, 19(2):24–26, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1963:TW

[Cha63b] Stuart Chase. Two worlds. Bulletin of the Atomic Scientists, 19 (6):18–20, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chamberlain:1964:LEM

[Cha64] Mary-Stuart Chamberlain. Letter to the Editor: Mythology of the atom. Bulletin of the Atomic Scientists, 20(1):26, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lil63].

Chalfont:1966:PP

[Cha66] Lord Chalfont. Prospects of peace. Bulletin of the Atomic Scientists, 22(5):2–4, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chai:1968:CCF

[Cha68a] Winberg Chai. Current comments: On Fulbright review. Bulletin of the Atomic Scientists, 24(1):64, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1968:VPB

[Cha68b] Sherret S. Chase. The vanishing plant breeder: Who will expand our food supply? Bulletin of the Atomic Scientists, 24(10):10– 13, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chang:1969:CSC

[Cha69a] Parris H. Chang. China’s scientists in the Cultural Revolution. Bulletin of the Atomic Scientists, 25(5):19–20, 40, May 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chapman:1969:CSP

[Cha69b] Richard L. Chapman. Congress and science policy: The organizational dilemma. Bulletin of the Atomic Scientists, 25(3):4–7, 28, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1969:AFS

[Cha69c] Sherret S. Chase. Anti-famine strategy: Genetic engineering for food. Bulletin of the Atomic Scientists, 25(8):2–6, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chandra:1970:CSS

[Cha70a] Helen Chandra. Chandrasekhar on scientists and society. Bulletin of the Atomic Scientists, 26(9):11–14, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1970:ACF

[Cha70b] Stuart Chase. Against a common fate. Bulletin of the Atomic Scientists, 26(5):9–10, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chau:1972:COL

[Cha72] W. Y. Chau. Communications: Official languages of U.N. Bulletin of the Atomic Scientists, 28(4):3, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bru72].

Chandrasekhar:1973:CHI

[Cha73a] S. Chandrasekhar. Copernicus — from humanism to inquisition. Bulletin of the Atomic Scientists, 29(6):27–30, June 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1973:CRC

[Cha73b] Stuart Chase. The Club of Rome and its computer. Bulletin of the Atomic Scientists, 29(3):36–39, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1974:BBG

[Cha74] Stuart Chase. Books: Toward Global Equilibrium: Collected Papers, edited by Dennis L. Meadows and Donella H. Meadows. Bulletin of the Atomic Scientists, 30(2):46–47, February 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chandrasekhar:1975:SFM

[Cha75a] S. Chandrasekhar. Of some famous men: Verifying the Theory of Relativity. Bulletin of the Atomic Scientists, 31(6):17–22, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1975:CCN

[Cha75b] Dennis J. Chase. A commentary: Clouding the nuclear reactor debate. Bulletin of the Atomic Scientists, 31(2):39–40, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Com75h].

Chase:1975:LC

[Cha75c] Stuart Chase. The last civilization. Bulletin of the Atomic Scientists, 31(1):37–40, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1975:OWN

[Cha75d] Stuart Chase. One world or none. Bulletin of the Atomic Scientists, 31(10):5–6, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1976:NNA

[Cha76] Stuart Chase. No nation alone. Bulletin of the Atomic Scientists, 32(10):42–43, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charney:1977:LCP

[Cha77] Stephen Charney. Letter: Cigarette pests. Bulletin of the Atomic Scientists, 33(7):4, 6, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chamberlain:1978:SPT

[Cha78a] Owen Chamberlain. Scientists protest trials in USSR. Bulletin of the Atomic Scientists, 34(8):9, 50–51, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1978:DQR

[Cha78b] Stuart Chase. Don Quixote rides again. Bulletin of the Atomic Scientists, 34(3):7, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sto78b].

Chandrasekhar:1979:BRB

[Cha79a] Subrahmanyan Chandrasekhar. Book reviews: The First Three Minutes, by Steven Weinberg. Bulletin of the Atomic Scientists, 35(3):64–67, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chase:1979:WB

[Cha79b] Stuart Chase. World without boundaries. Bulletin of the Atomic Scientists, 35(3):82–83, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chalk:1982:MC

[Cha82a] Rosemary Chalk. The miners’ canary. Bulletin of the Atomic Scientists, 38(2):16–22, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chalk:1982:SS

[Cha82b] Rosemary Chalk. Secrecy and science. Bulletin of the Atomic Scientists, 38(5):9–10, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chalk:1982:WNS

[Cha82c] Rosemary Chalk. Women and the national security debate. Bulletin of the Atomic Scientists, 38(7):44–46, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HC83].

Chalk:1983:SSC

[Cha83a] Rosemary Chalk. Security and scientific communication. Bulletin of the Atomic Scientists, 39(7):19–23, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Originally presented at the annual meeting of the American Association for the Advancement of Science in Denver in May 1983: see addendum [Ano83b].

Chandrasekhar:1983:CMB

[Cha83b] Subrahmanyan Chandrasekhar. Commentary: Moore’s “Nuclear Energy’. Bulletin of the Atomic Scientists, 39(2):54, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Report of a discussion between Moore and Chandrasekhar on the viewing of sculptures.

Chaiklin:1984:CWW

[Cha84] Seth Chaiklin. Commentary: On What would Vannevar Bush say? Bulletin of the Atomic Scientists, 40(6):51–52, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha84d].

Chappell:1985:USP

[Cha85a] David Chappell. U.S. security and the Philippines: an exchange. Bulletin of the Atomic Scientists, 41(5):38–40, May 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See rejoinder [Kes85b].

Charles:1985:WCN

[Cha85b] David Charles. Who controls NATO’s nuclear weapons? Bulletin of the Atomic Scientists, 41(4):45–48, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chalk:1986:CDS

[Cha86] Rosemary Chalk. Continuing debate over science and secrecy. Bulletin of the Atomic Scientists, 42(3):14–16, March 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charles:1987:NLA

[Cha87] Daniel Charles. NATO looks for arms control loopholes. Bulletin of the Atomic Scientists, 43(7):7–12, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chapman:1988:CW

[Cha88a] Gary Chapman. Computer wars. Bulletin of the Atomic Scientists, 44(9):14, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chapman:1988:MNH

[Cha88b] Gary Chapman. Move over, Nintendo, here comes SDI’s National Test Bed. Bulletin of the Atomic Scientists, 44(9):12–14, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charles:1988:PVC

[Cha88c] Daniel Charles. The people vs. the complex. Bulletin of the Atomic Scientists, 44(1):29–30, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chapman:1989:SRB

[Cha89a] Gary Chapman. Smart rocks, Brilliant Pebbles, genius dust? Bulletin of the Atomic Scientists, 45(9):10–11, 49, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charles:1989:ETW

[Cha89b] Dan Charles. Exporting trouble — West Germany’s freewheeling nuclear business. Bulletin of the Atomic Scientists, 45(3):21–27, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [The89a, Mül89].

Charles:1989:KSS

[Cha89c] Dan Charles. Keeping semiconductors safe for democracy. Bulletin of the Atomic Scientists, 45(9):8–10, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charles:1989:TA

[Cha89d] Dan Charles. The Transnuklear affair. Bulletin of the Atomic Scientists, 45(3):23, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chayes:2006:ANF

[Cha06a] Sarah Chayes. Afghanistan: The night fairies. Bulletin of the Atomic Scientists, 62(2):17–19, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/2/17.

Chayes:2006:SWA

[Cha06b] Sarah Chayes. Spinning the war in Afghanistan. Bulletin of the Atomic Scientists, 62(5):54–61, September/October 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/5/54.

Chang:2011:IFS

[Cha11] Soon Heung Chang. The implications of Fukushima: The South Korean perspective. Bulletin of the Atomic Scientists, 67(4): 18–22, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/4/18.

Chansoria:2016:BRA

[Cha16] Monika Chansoria. To ban or regulate autonomous weapons: An Indian response. Bulletin of the Atomic Scientists, 72(2):120– 121, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chew:1950:AFT

[Che50] Geoffrey Chew. Academic freedom on trial at the University of California. Bulletin of the Atomic Scientists, 6(11):333–336, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cherwell:1952:IWB

[Che52] Lord Cherwell. “It would have been worth millions”: Statement before the House of Lords, July 5, 1951 [on if Klaus Fuchs could have been prevented from giving atomic secrets to Russia]. Bulletin of the Atomic Scientists, 8(5):148, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chew:1956:PP

[Che56] Geoffrey Chew. Passport problems. Bulletin of the Atomic Scientists, 12(1):26–28, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cheever:1970:IOS

[Che70] Daniel S. Cheever. Inner and outer space: Marine science and ocean politics. Bulletin of the Atomic Scientists, 26(2):22, 29– 34, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cheh:1982:APS

[Che82] Mary M. Cheh. Atoms for peace: Secrecy: why is it still with us? Bulletin of the Atomic Scientists, 38(10):66–70, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chevrier:1988:BRB

[Che88] Marie Isabelle Chevrier. Book review: Gene Wars: Military Control over the New Genetic Technologies, by Charles Piller and Keith R. Yamamoto. Bulletin of the Atomic Scientists, 44(8): 47–48, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chernoff:1989:LBH

[Che89] Ralph Chernoff. Letter: Best hope is fewer bombs. Bulletin of the Atomic Scientists, 45(8):42, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano89-28].

Chepesiuk:1997:ST

[Che97] Ron Chepesiuk. A sea of trouble? Bulletin of the Atomic Scientists, 53(5):40–44, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kaf98].

Cheng:2020:EOE

[Che20] Yangyang Cheng. The edge of our existence. Bulletin of the Atomic Scientists, 76(6):315–320, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1846417.

Chisholm:1957:SS

[Chi57] Brock Chisholm. The search for security. Bulletin of the Atomic Scientists, 13(10):366–368, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chisholm:1959:BWD

[Chi59a] Brock Chisholm. Biological warfare: Demand for answers? Bulletin of the Atomic Scientists, 15(5):209–211, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chisholm:1959:LES

[Chi59b] Brock Chisholm. Letter to the Editor: Secret biological warfare possible? Bulletin of the Atomic Scientists, 15(10):401, 433, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chisholm:1959:RIT

[Chi59c] Brock Chisholm. Reduction of international tensions. Bulletin of the Atomic Scientists, 15(7):308–309, 314, September 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Presented, in a somewhat different form, at the Third Pugwash Conference.

Chinn:1969:ISC

[Chi69] Herman I. Chinn. International scientific co-operation. Bulletin of the Atomic Scientists, 25(9):34–35, 47, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Childress:1978:BRB

[Chi78] James Childress. Book reviews: Just and Unjust Wars: A Moral Argument with Historical Illustrations, by Michael Walzer. Bulletin of the Atomic Scientists, 34(8):44–48, October 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Childers:1982:CMA

[Chi82] Barry Childers. Commentary: More anti-MAD mythology. Bulletin of the Atomic Scientists, 38(2):47, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Buc81].

Chivvis:2017:HWR

[Chi17] Christopher S. Chivvis. Hybrid war: Russian contemporary political warfare. Bulletin of the Atomic Scientists, 73(5):316–321, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1362903.

Chou:1958:PPB

[Cho58] Pei-Yuan Chou. Population, production, and birth control. Bulletin of the Atomic Scientists, 14(8):325, 333, October 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Choriki:1981:REH

[Cho81] Danny Stephen Hifumi Choriki. 1980 Rabinowitch Essay on how to eliminate the threat of nuclear war: Begin with yourself. Bulletin of the Atomic Scientists, 37(4):41–44, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cho:1986:PAC

[Cho86a] Lee-Jay Cho. Population: Averting crisis in Asia? Bulletin of the Atomic Scientists, 42(4):30–33, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Choucri:1986:PDC

[Cho86b] Nazli Choucri. Population: Demographics and conflict. Bulletin of the Atomic Scientists, 42(4):24–25, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chopra:1995:BDB

[Cho95] Jarat Chopra. Back to the drawing board. Bulletin of the Atomic Scientists, 51(2):29–35, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Choffnes:2001:GL

[Cho01] Eileen Choffnes. Germs on the loose. Bulletin of the Atomic Scientists, 57(2):57–61, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/57.

Choffnes:2002:BNL

[Cho02] Eileen Choffnes. Bioweapons: New labs, more terror? Bulletin of the Atomic Scientists, 58(5):29–32, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/5/29.

Choudhury:2003:ACT

[Cho03] Upendra Choudhury. Article commentary: Too close for comfort. Bulletin of the Atomic Scientists, 59(2):22–25, March 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/2/22.

Caesar:1971:CMI

[CHP71] Irving Caesar, James H. Harger, and Frederik Pohl. Communications: Military–industrial complex. Bulletin of the Atomic Scientists, 27(6):2–3, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Christy:1968:BBA

[Chr68] Francis T. Christy, Jr. Books: Oceanography 1966: Achievement and Opportunities. Bulletin of the Atomic Scientists, 24(2):23– 24, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Christianson:1971:NTD

[Chr71] Gale Edward Christianson. Nuclear tyranny and the divine right of kings. Bulletin of the Atomic Scientists, 27(1):44–46, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Churchill:1950:WCE

[Chu50] Winston Churchill. Winston Churchill on European unity and settlement with Russia. Bulletin of the Atomic Scientists, 6(6): 168–170, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Churchill:1954:MBR

[Chu54a] Winston Churchill. [More on the British reaction]. Bulletin of the Atomic Scientists, 10(5):166–167, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Churchill:1954:SBH

[Chu54b] Winston Churchill. Speech before House of Commons, November 3, 1953. Bulletin of the Atomic Scientists, 10(1):11, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Churchill:1954:SHC

[Chu54c] Winston Churchill. Speech in House of Commons, December 17, 1953. Bulletin of the Atomic Scientists, 10(2):42, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Church:1960:PPG

[Chu60] Senator Frank Church. A proposal for progress at Geneva. Bulletin of the Atomic Scientists, 16(3):88, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Church:1976:CAS

[Chu76] Frank Church. Covert action: swampland of American foreign policy. Bulletin of the Atomic Scientists, 32(2):7–11, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Low76].

Chugaev:1993:KC

[Chu93] Sergei Chugaev. Khasbulatov & Co. Bulletin of the Atomic Scientists, 49(1):26–31, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chugonova:1994:BPR

[Chu94a] Nina Chugonova. Baltic pride, Russian tears. Bulletin of the Atomic Scientists, 50(5):22–29, September/October 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chugonova:1994:CNZ

[Chu94b] Nina Chugonova. Cosmonauts number zero. Bulletin of the Atomic Scientists, 50(3):16–21, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chuen:2003:BW

[Chu03] Cristina Chuen. Beneath the waves. Bulletin of the Atomic Scientists, 59(3):71–72, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/3/71.

Chuen:2004:URE

[Chu04] Cristina Chuen. Unfreezing Russia’s economy. Bulletin of the Atomic Scientists, 60(3):65–66, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/65.

Chuen:2005:DSS

[Chu05] Cristina Chuen. Don’t sweat the suitcase. Bulletin of the Atomic Scientists, 61(1):69–70, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/69.

Cicarelli:1974:WHT

[Cic74] James Cicarelli. Whatever happened to the turbine car? A revolutionary break-through thwarted by auto industry inertia. Bulletin of the Atomic Scientists, 30(10):24–29, December 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SC75].

Cimbala:1984:MMP

[Cim84] Stephen J. Cimbala. Midgetman: major problems. Bulletin of the Atomic Scientists, 40(2):7–8, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cirincione:1998:RF

[Cir98] Joseph Cirincione. Rush to failure. Bulletin of the Atomic Scientists, 54(3):23–25, 68, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cirincione:2000:RDI

[Cir00] Joseph Cirincione. Republicans do it better. Bulletin of the Atomic Scientists, 56(5):17–19, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/17.

Cirincione:2005:LL

[Cir05a] Joseph Cirincione. Lessons lost. Bulletin of the Atomic Scientists, 61(6):42–53, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/6/42.

Cirincione:2005:OCW

[Cir05b] Joseph Cirincione. Not one claim was true. Bulletin of the Atomic Scientists, 61(1):65–67, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/65.

Cirincione:2007:ACN

[Cir07a] Joseph Cirincione. Arms Control’s new moment. Bulletin of the Atomic Scientists, 63(4):21, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/21.

Cirincione:2007:CC

[Cir07b] Joseph Cirincione. Change of course. Bulletin of the Atomic Scientists, 63(2):20, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/2/20.

Cisler:1952:EPS

[Cis52] Walker L. Cisler. Electric power systems and nuclear power. Bulletin of the Atomic Scientists, 8(8):279–282, November 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Civiak:2002:NS

[Civ02] Robert Civiak. The need for speed. Bulletin of the Atomic Scientists, 58(4):38–43, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/38.

Chuen:2001:RBW

[CJ01] Cristina Chuen and Michael Jasinski. Russia’s blue water blues. Bulletin of the Atomic Scientists, 57(1):65–69, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/1/65.

Corkhill:2024:SAN

[CJLT24] Claire Corkhill, Malcolm Joyce, Derek Lacey, and Nigel Thrift. Small and advanced nuclear reactors: Closing the fuel cycle? Bulletin of the Atomic Scientists, 81(1):43–47, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2441046.

Conetta:1998:IT

[CK98] Carl Conetta and Charles Knight. Inventing threats. Bulletin of the Atomic Scientists, 54(2):32–38, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cross:2020:TFC

[CK20] Glenn Cross and Lynn Klotz. Twenty-first century perspectives on the Biological Weapon Convention: Continued relevance or toothless paper tiger. Bulletin of the Atomic Scientists, 76(4):185–191, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1778365.

Chavarria:2010:HCW

[CKA10] Gabriela Chavarria, Kim Knowlton, and Dylan Atchley. The human-climate-wildlife nexus. Bulletin of the Atomic Scientists, 66(1):48–56, January 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/1/48.

Chugonova:1994:TOW

[CKN^{+}94] Nina Chugonova, Yevgeni Kiriushin, Sergei Nefiodov, Victor Volkov, and Yuri Savochkin. In their own words. Bulletin of the Atomic Scientists, 50(3):22–23, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cortright:2000:LC

[CL00] David Cortright and George A. Lopez. The limits of coercion. Bulletin of the Atomic Scientists, 56(6):18–20, November/ December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
6/18.

Carter:2006:FCF

[CL06] Ashton B. Carter and Stephen A. Lamontagne. A fuel-cycle fix. Bulletin of the Atomic Scientists, 62(1):24–25, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/1/24.

Claiborne:1954:LEB

[Cla54] Robert W. Claiborne. Letter to the Editor: Offensive or Defensive Weapons? Bulletin of the Atomic Scientists, 10(1):32, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clark:1956:GRE

[Cla56] Stanley H. Clark. Genetic radiation exposures in the field of medicine. Bulletin of the Atomic Scientists, 12(1):14–18, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clausen:1959:LER

[Cla59] Harold C. Clausen. Letter to the Editor: Religion without revelation? Bulletin of the Atomic Scientists, 15(10):433, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor59a].

Clark:1961:CTE

[Cla61] W. H. Clark. Chemical and thermonuclear explosives. Bulletin of the Atomic Scientists, 17(9):356–360, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clark:1963:CD

[Cla63] Joseph S. Clark. Congress and disarmament. Bulletin of the Atomic Scientists, 19(7):3–8, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GG63].

Clark:1966:RSC

[Cla66] Joseph S. Clark. Reports: Senator Clark on the White House Conference. Bulletin of the Atomic Scientists, 22(3):30–33, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Classen:1968:FP

[Cla68] H. George Classen. Fact and purpose. Bulletin of the Atomic Scientists, 24(3):36–38, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clarke:1974:UDA

[Cla74] Duncan L. Clarke. Ups and downs of arms control: Beset by bureaucratic shuffles, U.S. agency faces an uncertain future. Bulletin of the Atomic Scientists, 30(7):44–49, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clayton:1977:PDA

[Cla77] Bruce Douglas Clayton. Planning for the day after doomsday: abandoning hope that nuclear war can be prevented may be the only hope for prevention. Bulletin of the Atomic Scientists, 33 (7):49–53, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mue77, Lev77b, Pal77b].

Clarke:1981:ACF

[Cla81] Duncan L. Clarke. Arms control and foreign policy under Reagan. Bulletin of the Atomic Scientists, 37(9):12–19, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Buc82].

Clark:1982:EE

[Cla82] Edwin H. Clark II. Election 1982: The environment. Bulletin of the Atomic Scientists, 38(8):14–15, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Claghorn:1990:LLD

[Cla90] Marguerite S. Claghorn. Letter: … but did better in other subjects. Bulletin of the Atomic Scientists, 46(7):47, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clausen:1991:RSS

[Cla91] Peter Clausen. Report: SDI: Star warriors try again. Bulletin of the Atomic Scientists, 47(5):9–10, 42, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clarke:1992:WD

[Cla92] Arthur C. Clarke. What is to be done? Bulletin of the Atomic Scientists, 48(4):10–13, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clark:2016:UCF

[Cla16] Bryan Clark. Undersea cables and the future of submarine competition. Bulletin of the Atomic Scientists, 72(4):234–??, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clarke:2020:WD

[Cla20]  Arthur C. Clarke. 1992: What is to be done?  Bul-

letin of the Atomic Scientists, 76(6):391–394, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847498.

Cahn:1988:NCS

[CLD88] Anne H. Cahn, Martha C. Little, and Stephen Daggett. Nunn and contractors sell ALPS. Bulletin of the Atomic Scientists, 44 (5):10–12, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sch88a].

CSCFAS:1948:LCP

[Cle48a] Committee on Secrecy and Federation of American Scientists Clearance. Loyalty clearance procedures in research laboratories. Bulletin of the Atomic Scientists, 4(4):111–114, April 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

CSCFAS:1948:SIC

[Cle48b] Committee on Secrecy and Federation of American Scientists Clearance. Some individual cases of clearance procedures. Bulletin of the Atomic Scientists, 4(9):281–285, September 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

CSCFAS:1948:HFS

[Cle48c] Committee on Secrecy and Federation of Atomic Scientists Clearance. How far should military censorship extend? Bulletin of the Atomic Scientists, 4(6):163–165, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clemens:1967:OSS

[Cle67] Walter C. Clemens, Jr. Outer space, strategy, and arms control. Bulletin of the Atomic Scientists, 23(9):24–28, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clemens:1970:EW

[Cle70] Walter C. Clemens, Jr. The ecology of weaponry. Bulletin of the Atomic Scientists, 26(7):27–31, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clemens:1971:SMO

[Cle71] Walter C. Clemens, Jr. Sakharov: A man for our times. Bulletin of the Atomic Scientists, 27(10):4–6, 51–56, December 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clemens:1974:BBP

[Cle74] Walter C. Clemens, Jr. Books: The Politics of Nuclear Proliferation, by George H Quester. Bulletin of the Atomic Scientists, 30 (7):63–64, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clem:1982:SUE

[Cle82] Ralph S. Clem. The Soviet Union: Ethnicity and its implications. Bulletin of the Atomic Scientists, 38(6):53–58, June 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clements:1987:IVN

[Cle87a] Kevin P. Clements. International views: New Zealand’s antinuclear stand. Bulletin of the Atomic Scientists, 43(2):32–34, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clements:1987:NZP

[Cle87b] Kevin P. Clements. New Zealand paying for nuclear ban. Bulletin of the Atomic Scientists, 43(6):41–44, July/August 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clements:1989:WTB

[Cle89] Kevin P. Clements. Will test ban conference self-destruct? Bulletin of the Atomic Scientists, 45(6):16–18, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tov89].

Clements:1992:RNZ

[Cle92] Kevin P. Clements. Report: New Zealand: Kiwi no-nuke policy at risk. Bulletin of the Atomic Scientists, 48(1):7–8, 44, January/ February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clemens:1995:JBW

[Cle95] Walter C. Clemens, Jr. Jerome B. Wiesner, 1915–1994. Bulletin of the Atomic Scientists, 51(1):10, January/February 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clegg:1999:NTA

[Cle99] Liz Clegg. NGOs take aim. Bulletin of the Atomic Scientists, 55(1):49–51, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/1/49.

Clemens:2001:MDW

[Cle01] Walter C. Clemens, Jr. Missile defense: Who terminates a treaty? Bulletin of the Atomic Scientists, 57(6):38–43, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/57/6/38.

Clearwater:2004:FOG

[Cle04a] John M. Clearwater. The first one to get away. Bulletin of the Atomic Scientists, 60(6):22–27, November/December 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/6/22.

Clemens:2004:ABS

[Cle04b] Walter C. Clemens. Almost back to square one. Bulletin of the Atomic Scientists, 60(5):22–25, September/October 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/5/22.

Clemente:2007:SH

[Cle07] Jonathan D. Clemente. In sickness and in health. Bulletin of the Atomic Scientists, 63(2):??, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/2/38.

Cleetus:2011:FCG

[Cle11] Rachel Cleetus. Finding common ground in the debate between carbon tax and cap-and-trade policies. Bulletin of the Atomic Scientists, 67(1):19–27, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/1/19.

Clinton:1977:NDC

[Cli77] Richard Lee Clinton. The never-to-be-developed countries of Latin America. Bulletin of the Atomic Scientists, 33(8):19–26, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clotfelter:1971:NNI

[Clo71] James Clotfelter. Notes on the new isolationists. Bulletin of the Atomic Scientists, 27(2):37, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cahn:1974:SNH

[CLR74] Anne H. Cahn, F. A. Long, and George Rathjens. The search for a new handle on arms control. Bulletin of the Atomic Scientists, 30(4):7, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Chayes:1975:TTS

[CLR75] Abram Chayes, Franklin A. Long, and George W. Rathjens. Threshold treaty: A step backward. Bulletin of the Atomic Scientists, 31(1):16, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Choffnes:2006:BNW

[CLR06] Eileen R. Choffnes, Stanley M. Lemon, and David A. Relman. A brave new world in the life sciences. Bulletin of the Atomic Scientists, 62(5):26–33, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/26.

Cassel:1984:M

[CM84] Christine K. Cassel and Michael McCally. Medicine. Bulletin of the Atomic Scientists, 40(10):16s–17s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1991:PIR

[CM91] Avner Cohen and Marvin Miller. Perspective: Iraq and the rules of the nuclear game. Bulletin of the Atomic Scientists, 47(6):10– 11, 43, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:2012:CCS

[CM12] Steven Cohen and Alison Miller. Climate change 2011: a status report on US policy. Bulletin of the Atomic Scientists, 68(1):39– 49, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/1/39.

Coster-Mullen:1995:LA

[CMM95] John Coster-Mullen and Mike Moore. Letters: Also the aggressors. Bulletin of the Atomic Scientists, 51(4):76, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo95h].

Cooney:1962:LER

[CMMF62] Thomas E. Cooney, William J. Mechan, C. R. May, and Lloyd Fallers. Letter to the Editor: Readers score Fallers on Snow. Bulletin of the Atomic Scientists, 18(2):31–32, February 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fal61a].

Cohen:1979:CCL

[CMR79] Bernard L. Cohen, Karl Z. Morgan, and Joseph Rotblat. Commentary: On cancer and low level radiation: What is the misunderstanding all about? Bulletin of the Atomic Scientists, 35 (2):53–59, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor78, Rot78b].

Christie:1988:SRN

[CN88] Daniel J. Christie and Linden Nelson. Student reactions to nuclear education. Bulletin of the Atomic Scientists, 44(6):22–23, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bye88].

Cochran:1990:SNW

[CN90] Thomas B. Cochran and Robert S. (Robert Stan) Norris. Soviet nuclear warhead production, volume NWD 90-3 of Nuclear weapons databook working papers. Natural Resources Defense Council, Washington, DC, USA, 1990. 32 pp. LCCN U264.5.S65.C67 1990.

Cochran:1991:FLS

[CN91] Thomas B. Cochran and Robert S. Norris. A first look at the Soviet bomb complex. Bulletin of the Atomic Scientists, 47(4):

25–31, May 1991. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://docs.nrdc.org/nuclear/
nuc_05009101a_101.pdf.

Cochran:1995:MRB

[CNB95] Thomas B. Cochran, Robert S. (Robert Stan) Norris, and Oleg Bukharin. Making the Russian bomb: from Stalin to Yeltsin. Westview Press, Boulder, CO, USA, 1995. ISBN 0-8133-2328-2 (hardcover). LCCN U264.S65 C63 1995.

Clemens:1974:LEH

[CO74] Walter C. Clemens, Jr. and Deborah Oakley. Letter: Ersatz hot dog. Bulletin of the Atomic Scientists, 30(7):2, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kru74].

Cooper:1985:LSC

[CO85] Robert S. Cooper and Severo M. Ornstein. Letter: Strategic computing initiative: an exchange. Bulletin of the Atomic Scientists, 41(1):54–56, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [OSS84] and comment [Der85].

Clearwater:2003:LC

[CO03] John Clearwater and David O’Brien. O lucky Canada. Bulletin of the Atomic Scientists, 59(4):60–65, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/4/60.

Cohen:2006:LSS

[CO06] Roberta Cohen and William G. O’Neill. Last stand in Sudan. Bulletin of the Atomic Scientists, 62(2):51–58, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/2/51.

Clingerman:2014:PGW

[CO14] Forrest Clingerman and Kevin J. O’Brien. Playing God: Why religion belongs in the climate engineering debate. Bulletin of the Atomic Scientists, 70(3):27–37, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coale:1947:RVA

[Coa47] Ansley Coale. Reducing vulnerability to atomic attack … protective methods that may reduce the potential consequences of an attack by atomic bombs. Bulletin of the Atomic Scientists, 3 (3):71–74, 98, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coatsworth:1984:ACA

[Coa84] John H. Coatsworth. Agenda 1984: Central America. Bulletin of the Atomic Scientists, 40(1):10–12, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coad:1988:PST

[Coa88] Thomas J. Coad. Perspective: Scientists: be there or b^{2}. Bulletin of the Atomic Scientists, 44(5):5–6, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Eps88a].

Cobb:1978:LPP

[Cob78] John C. Cobb. Letter: Down a primrose path? Bulletin of the Atomic Scientists, 34(5):4–5, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coburn:1988:LDD

[Cob88] Patricia Coburn. Letter: Dangerous, destabilizing D-5. Bulletin of the Atomic Scientists, 44(3):56, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa88b].

Cockcroft:1950:DFN

[Coc50] Sir John Cockcroft. The development and future of nuclear energy. Bulletin of the Atomic Scientists, 6(11):325–331, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cockroft:1951:RBA

[Coc51] Sir John Cockroft. Report on the British Atomic Energy Research Establishment: An editorial note. Bulletin of the Atomic Scientists, 7(7–8):229, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cockcroft:1952:PS

[Coc52] Sir John Cockcroft. Pure science. Bulletin of the Atomic Scientists, 8(4):101–103, 110, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cockcroft:1955:FAE

[Coc55] Sir John Cockcroft. The future of atomic energy. Bulletin of the Atomic Scientists, 11(8):285–288, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cockcroft:1959:PUA

[Coc59] Sir John Cockcroft. Peaceful uses of atomic energy: The Second International Conference. Bulletin of the Atomic Scientists, 15 (1):18–21, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cochran:1981:CSC

[Coc81a] Thomas B. Cochran. Commentary: Soviet ‘civil’ reactors. Bulletin of the Atomic Scientists, 37(5):63, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar81b].

Cochran:1981:SNP

[Coc81b] Thomas B. Cochran. Secrecy and nuclear power. Bulletin of the Atomic Scientists, 37(7):37–41, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cochran:1989:BSE

[Coc89] Thomas B. Cochran. Black Sea experiment only a START. Bulletin of the Atomic Scientists, 45(9):12–16, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hun90].

Coe:1998:LNE

[Coe98] Steven Coe. Letter: NATO expansion provocative. Bulletin of the Atomic Scientists, 54(3):67, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffin:1962:ECTa

[Cof62a] Tristram Coffin. Events and comments: Tristram Coffin reports. Bulletin of the Atomic Scientists, 18(3):32–33, March 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffin:1962:ECTb

[Cof62b] Tristram Coffin. Events and comments: Tristram Coffin reports. Bulletin of the Atomic Scientists, 18(4):31–33, April 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffin:1962:ECTc

[Cof62c] Tristram Coffin. Events and comments: Tristram Coffin reports. Bulletin of the Atomic Scientists, 18(5):42–43, May 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffin:1962:ECTd

[Cof62d] Tristram Coffin. Events and comments: Tristram Coffin reports. Bulletin of the Atomic Scientists, 18(6):44–45, June 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffin:1962:PP

[Cof62e] Tristram Coffin. Peace in politics. Bulletin of the Atomic Scientists, 18(8):43–44, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffey:1965:FCB

[Cof65a] J. I. Coffey. Forum: The Chinese and ballistic missile defense. Bulletin of the Atomic Scientists, 21(10):17–19, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffey:1965:FTN

[Cof65b] Rosemary Klineberg Coffey. Forum: Three on nuclear policies: I. The heart of deterrence. Bulletin of the Atomic Scientists, 21(4): 27–29, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffey:1966:BBB

[Cof66] J. I. Coffey. Books: China and the Bomb, by Morton H. Halperin. Bulletin of the Atomic Scientists, 22(5):35–36, May 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffin:1967:CCL

[Cof67] Tristram Coffin. Comments: Congress: Its lost sacred powers. Bulletin of the Atomic Scientists, 23(10):35–37, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffey:1970:SAA

[Cof70] J. I. Coffey. The Soviet ABM and arms control. Bulletin of the Atomic Scientists, 26(1):39–43, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coffey:1973:SS

[Cof73] J. I. Coffey. The savor of SALT. Bulletin of the Atomic Scientists, 29(5):9–15, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coggle:1982:BRB

[Cog82] John E. Coggle. Book review: Hiroshima and Nagasaki — the Physical, Medical, and Social Effects of the Atomic Bombings, by The Committee for the Compilation of Materials on Damage Caused by the Atomic Bombs in Hiroshima and Nagasaki, translated by Eisei Ishikawa and David L. Swain. Bulletin of the Atomic Scientists, 38(2):33–34, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1948:REM

[Coh48] Karl Cohen. A re-examination of the McMahon Act. Bulletin of the Atomic Scientists, 4(1):7–10, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1953:APR

[Coh53a] Karl Cohen. Atomic power as a risk venture. Bulletin of the Atomic Scientists, 9(8):305–308, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohn:1953:TRG

[Coh53b] Victor Cohn. “Treat reporters as gentlemen”. Bulletin of the Atomic Scientists, 9(9):329–330, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1960:LEC

[Coh60] Mortimer T. Cohen. Letter to the Editor: Chances of war. Bulletin of the Atomic Scientists, 16(10):423, December 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra60].

Cohen:1962:LEW

[Coh62] Sidney Cohen. Letter to the Editor: Weapons criteria. Bulletin of the Atomic Scientists, 18(3):39, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1969:BRB

[Coh69] Marshall H. Cohen. Book review: The birth of a large telescope: The Story of Jodrell Bank, by Sir Bernard Lovell. Bulletin of the Atomic Scientists, 25(6):40, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1972:UUD

[Coh72] Benjamin I. Cohen. U.S.–USSR in developing countries. Bulletin of the Atomic Scientists, 28(2):22–28, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1974:OVP

[Coh74] Bernard L. Cohen. An opposing view: Perspectives on the nuclear debate. Bulletin of the Atomic Scientists, 30(8):35–39, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1975:LFI

[Coh75a] Bernard I. Cohen. Letters: False impressions. Bulletin of the Atomic Scientists, 31(1):4, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day74d, Hol74].

Cohen:1975:LM

[Coh75b] Bernard L. Cohen. Letters: Not so malevolent. Bulletin of the Atomic Scientists, 31(2):2–3, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DeN74b, DeN75].

Cohen:1976:LER

[Coh76a] Bernard L. Cohen. Letter: Effects of radon. Bulletin of the Atomic Scientists, 32(5):3–4, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com76a].

Cohen:1976:LED

[Coh76b] Bernard L. Cohen. Letter: Exaggerated danger. Bulletin of the Atomic Scientists, 32(10):6, 8, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds76].

Cohen:1976:PT

[Coh76c] Bernard L. Cohen. The potentialities of terrorism. Bulletin of the Atomic Scientists, 32(6):34–35, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1976:REI

[Coh76d] Bernard L. Cohen. Rebuttal: Environmental impacts of nuclear power due to radon emissions. Bulletin of the Atomic Scientists, 32(2):61–63, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com75e, Com76a].

Cohen:1979:MLL

[Coh79] Bernard L. Cohen. More on low-level ionizing radiation. Bulletin of the Atomic Scientists, 35(10):56–57, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor78, Rot78b].

Cohen:1981:HCU

[Coh81] Karl Cohen. Harold C. Urey 1893–1981. Bulletin of the Atomic Scientists, 37(5):8, 54–56, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1986:SAS

[Coh86] Danny Cohen. From an SDI advocate: “Software is within reach of current technology”. Bulletin of the Atomic Scientists, 42(2): 34, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohn:1987:SEG

[Coh87] Carol Cohn. Slick ’ems, glick ’ems, Christmas trees, and cookie cutters: Nuclear language and how we learned to pat the bomb. Bulletin of the Atomic Scientists, 43(5):17–24, June 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BB87, SHT^{+}87].

Cohen:1992:USS

[Coh92] Stephen P. Cohen. U.S. security in a separatist season. Bulletin of the Atomic Scientists, 48(6):28–32, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1993:SM

[Coh93a] Avner Cohen. A sacred matter. Bulletin of the Atomic Scientists, 49(5):39–41, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Say93, Day93].

Cohn:1993:DNN

[Coh93b] Avner Cohn. Did nukes nudge the PLO? Bulletin of the Atomic Scientists, 49(10):11–13, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1995:LEG

[Coh95a] Avner Cohen. Letter: The editors got it wrong. Bulletin of the Atomic Scientists, 51(2):69–70, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coh95b].

Cohen:1995:NMF

[Coh95b] Avner Cohen. Nonproliferation: Most favored nation. Bulletin of the Atomic Scientists, 51(1):44–53, January/February 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See objection to title [Coh95a].

Cohen:1996:PPP

[Coh96] Avner Cohen. Perspective: Peres: Peacemaker, nuclear pioneer. Bulletin of the Atomic Scientists, 52(3):16–17, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1998:TTW

[Coh98] Avner Cohen. And then there was one. Bulletin of the Atomic Scientists, 54(5):51–55, September/October 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:2000:BN

[Coh00] Avner Cohen. The bomb that never is. Bulletin of the Atomic Scientists, 56(3):22–23, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/56/3/22.

Cohen:2002:NRN

[Coh02] Roberta Cohen. Nowhere to run, no place to hide. Bulletin of the Atomic Scientists, 58(6):36–45, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/6/36.

Cole:1954:CBB

[Col54] Sterling Cole. Comment on the Hydrogen Bomb book. Bulletin of the Atomic Scientists, 10(9):368, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cole:1959:IAE

[Col59] Sterling Cole. International Atomic Energy Agency: The administrator’s view. Bulletin of the Atomic Scientists, 15(4):160–162, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cole:1974:EDP

[Col74] Henry S. Cole. An ‘environmental degradation preserve’: Let’s have a national monument to untrammeled economic growth. Bulletin of the Atomic Scientists, 30(7):53–55, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cole:1979:MTM

[Col79a] Richard W. Cole. More on Three Mile Island. Bulletin of the Atomic Scientists, 35(7):63, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lan79].

Colgate:1979:CCO

[Col79b] Stirling A. Colgate. Commentary: The case for one nuclear waste facility. Bulletin of the Atomic Scientists, 35(7):55–56, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Collins:1979:HNW

[Col79c] A. S. Collins, Jr. How a nuclear war would be fought … on land. Bulletin of the Atomic Scientists, 35(5):28–30, May 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Colaianni:1984:CEU

[Col84a] Ed Colaianni. Commentary: On ‘The economics of U.S. foreign aid’. Bulletin of the Atomic Scientists, 40(3):62, March 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch83].

Cole:1984:BRB

[Col84b] Leonard A. Cole. Book review: The New Politics of Science, by David Dickson. Bulletin of the Atomic Scientists, 40(9):48– 49, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cole:1984:YRY

[Col84c] Leonard A. Cole. Yellow rain or yellow journalism? Bulletin of the Atomic Scientists, 40(7):36–38, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Colton:1988:NHB

[Col88] Timothy J. Colton. No holds barred at Soviet conference. Bulletin of the Atomic Scientists, 44(8):7–9, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Colby:1993:PFG

[Col93a] Gale Colby. Perspective: Fabricating guilt. Bulletin of the Atomic Scientists, 49(8):12–13, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Collina:1993:BCB

[Col93b] Tom Zamora Collina. Bulletin: China bucks ban with bang. Bulletin of the Atomic Scientists, 49(10):3–4, December 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cole:1994:CAM

[Col94] Leonard A. Cole. Clearing the air in Minneapolis. Bulletin of the Atomic Scientists, 50(9):5–6, November 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

www.highbeam.com/doc/1G1-15879963.html.

Colgate:1995:LLV

[Col95] Rosie Colgate. Letter: The long view. Bulletin of the Atomic Scientists, 51(5):59–60, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol95c].

Cole:2000:CTD

[Col00] Leonard A. Cole. CBW terrorism, deconstructed. Bulletin of the Atomic Scientists, 56(5):58–59, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/58.

Cole:2003:GTH

[Col03] Leonard A. Cole. Gone today, here tomorrow? Bulletin of the Atomic Scientists, 59(5):66–69, September/October 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/5/66.

Coleman:2006:CNC

[Col06] David Coleman. Camelot’s nuclear conscience. Bulletin of the Atomic Scientists, 62(3):40–45, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/3/40.

Compton:1946:IAB

[Com46] Karl T. Compton. If the atomic bomb had not been used. Atlantic Monthly, ??(??):??, December 1946. ISSN 1072-7825 (print), 2151-9463 (electronic). URL http:/

/www.theatlantic.com/magazine/archive/1946/12/if-the-
atomic-bomb-had-not-been-used/376238/.

CFWC:1948:PDW

[Com48a] Committee to Frame a World Constitution. Preliminary draft of a World Constitution. Bulletin of the Atomic Scientists, 4(5): 145–150, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See explanation [Bor48] and comment [Tel48a].

Compton:1948:MJC

[Com48b] Arthur H. Compton. In memoriam: Joyce Clennam Stearns. Bulletin of the Atomic Scientists, 4(8):235, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Compton:1948:SS

[Com48c] Karl T. Compton. Science and security. Bulletin of the Atomic Scientists, 4(12):373–376, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Compton:1950:LPD

[Com50] Arthur H. Compton. Let the people decide! Bulletin of the Atomic Scientists, 6(3):74–75, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

CFCCCA:1951:CCW

[Com51] Commission of the Federal Council of the Churches of Christ in America. The Christian conscience and weapons of mass destruction. Bulletin of the Atomic Scientists, 7(4):115–117, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Compton:1953:BAP

[Com53] Arthur H. Compton. The birth of atomic power. Bulletin of the Atomic Scientists, 9(1):10–12, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Commager:1954:CCG

[Com54] Henry Steele Commager. Current comment: Guilt — and innocence — by association. Bulletin of the Atomic Scientists, 10(3): 111, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Committee:1957:PCH

[Com57a] Committee I. The Pugwash Conference: Hazards arising from the use of atomic energy in peace and war. Bulletin of the Atomic Scientists, 13(7):251, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Committee:1957:PCP

[Com57b] Committee II. The Pugwash Conference: Problems of nuclear weapon control. Bulletin of the Atomic Scientists, 13(7):252, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Committee:1957:PCS

[Com57c] Committee III. The Pugwash Conference: Social responsibility of scientists. Bulletin of the Atomic Scientists, 13(7):252, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Comar:1963:RFA

[Com63a] C. L. Comar. Radiation and food: Is action needed? Bulletin of the Atomic Scientists, 19(3):37–40, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Commoner:1963:SS

[Com63b] Barry Commoner. Scientific statesmanship. Bulletin of the Atomic Scientists, 19(8):6–10, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Commoner:1972:BDR

[Com72] Barry Commoner. A Bulletin dialogue: Response on “The Closing Circle”. Bulletin of the Atomic Scientists, 28(5):17, 42–56, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [EH72a, EH72b, MW72, Whi72, But73].

Comar:1973:BR

[Com73] C. L. Comar. The BEIR report. Bulletin of the Atomic Scientists, 29(7):2–3, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Comey:1974:WIC

[Com74] David Dinsmore Comey. Will idle capacity kill nuclear power? Uncle Sam may have to bail out some giants of American industry. Bulletin of the Atomic Scientists, 30(9):23–28, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Hur75, ML75, Com75f, Net75].

Comey:1975:ANP

[Com75a] David Dinsmore Comey. Advice to the nuclear power industry: The truth, nothing but the truth. Bulletin of the Atomic Scientists, 31(6):11, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [AC76].

Comey:1975:DGG

[Com75b] David Dinsmore Comey. Do not go gentle into that radiation zone. Bulletin of the Atomic Scientists, 31(9):45–47, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [vHC76].

Comey:1975:ES

[Com75c] David Dinsmore Comey. Elastic statistics. Bulletin of the Atomic Scientists, 31(5):4, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Comey:1975:FL

[Com75d] David Dinsmore Comey. Following the leader? Bulletin of the Atomic Scientists, 31(8):45, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Comey:1975:LUT

[Com75e] David Dinsmore Comey. Legacy of uranium tailings. Bulletin of the Atomic Scientists, 31(7):42–45, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Coh76d, SC76] and reply [Com76a].

Comey:1975:CC

[Com75f] David Dinsmore Comey. On cooking curves. Bulletin of the Atomic Scientists, 31(8):40–42, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ML75].

Comey:1975:PVT

[Com75g] David Dinsmore Comey. Points vs. trends. Bulletin of the Atomic Scientists, 31(8):44–45, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com74, Com75h].

Comey:1975:RCF

[Com75h] David Dinsmore Comey. Response: Chasing down the facts. Bulletin of the Atomic Scientists, 31(2):40–42, February 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha75b, ML75, Com75f, Net75].

Comey:1976:LBW

[Com76a] David Dinsmore Comey. Letter: Blowing in the wind. Bulletin of the Atomic Scientists, 32(2):63–65, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coh76a, Com75e, Coh76d].

Comey:1976:PTH

[Com76b] David Dinsmore Comey. The perfect Trojan horse. Bulletin of the Atomic Scientists, 32(6):33–34, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [RD76, Bir76].

CH:1981:PHA

[Com81] Commission on the Humanities. Perspective: “The humanities in American life”. Bulletin of the Atomic Scientists, 37(4):5–6, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Combs:2008:UM

[Com08] Jeff Combs. Uranium markets. Bulletin of the Atomic Scientists, 64(4):48–51, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/4/48.

Condon:1946:AR

[Con46a] Edward U. Condon. An appeal to reason. Bulletin of the Atomic Scientists, 1(7):6–7, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1946:SIC

[Con46b] Edward U. Condon. Science and international co-operation. Bulletin of the Atomic Scientists, 1(11):8–11, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1948:SGD

[Con48] Edward U. Condon. Scientists and the government: Dr. Condon writes to Senator Hickenlooper. Bulletin of the Atomic Scientists, 4(4):97–98, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1949:RG

[Con49] Edward U. Condon. Reflections on government. Bulletin of the Atomic Scientists, 5(6–7):179–181, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1952:SFG

[Con52] Edward U. Condon. Scientists and the Federal Government. Bulletin of the Atomic Scientists, 8(6):179–182, August 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1955:ECS

[Con55] Edward Condon. Edward Condon’s security clearance: Statement by Dr. Condon. Bulletin of the Atomic Scientists, 11(2): 66, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1958:TSB

[Con58] Edward U. Condon. Time to stop baiting scientists. Bulletin of the Atomic Scientists, 14(2):80–82, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Conrad:1960:LEB

[Con60] William A. Conrad. Letter to the Editor: Better the ban than the bomb. Bulletin of the Atomic Scientists, 16(3):113, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1964:FEW

[Con64] Edward U. Condon. Forum: Education for world understanding. Bulletin of the Atomic Scientists, 20(3):18–19, March 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Consolazio:1965:FDA

[Con65] William V. Consolazio. The fiscal dilemma of academic science. Bulletin of the Atomic Scientists, 21(2):15–18, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1967:BBI

[Con67] Edward U. Condon. Books: The Idea of a World University, by Michael Zweig. Bulletin of the Atomic Scientists, 23(8):34– 35, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Condon:1969:ULL

[Con69] Edward U. Condon. UFOs I have loved and lost. Bulletin of the Atomic Scientists, 25(10):6–8, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Connell:1981:BRB

[Con81a] Elizabeth B. Connell. Book review: The Politics of Contraception, by Carl Djerassi. Bulletin of the Atomic Scientists, 37(3):46, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Constable:1981:SPA

[Con81b] John D. Constable. Surgical problems among survivors. Bulletin of the Atomic Scientists, 37(5):22–25, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kle81].

Conrad:1986:SAC

[Con86] Thomas Conrad. South Africa circumvents embargo. Bulletin of the Atomic Scientists, 42(3):8–13, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Conrad:1988:BRB

[Con88] Thomas Conrad. Book review: Cross and Commissar: The Politics of Religion in Eastern Europe and the U.S.S.R., by Pedro Ramet. Bulletin of the Atomic Scientists, 44(4):50–51, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Conable:1990:RTW

[Con90a] Barbara B. Conable. For the record: Third World must overcome 1980s. Bulletin of the Atomic Scientists, 46(1):11–12, January/ February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Connor:1990:RHN

[Con90b] Tim Connor. Radiation and health: Nuclear workers at risk. Bulletin of the Atomic Scientists, 46(7):24–28, September 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:1956:PIS

[Coo56] Robert C. Cook. Part IV: Science and world economy: The population bomb. Bulletin of the Atomic Scientists, 12(8):295–298, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:1958:LEN

[Coo58] Barbara B. Cook. Letter to the Editor: No more defeatism. Bulletin of the Atomic Scientists, 14(5):201, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooper:1962:NWR

[Coo62a] Ralph S. Cooper. New world revisited. Bulletin of the Atomic Scientists, 18(4):38–39, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooper:1962:RP

[Coo62b] Ralph S. Cooper. Rocket propulsion. Bulletin of the Atomic Scientists, 18(3):22–28, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:1971:MRT

[Coo71] Robert E. Cook. The mist that rolled into the trenches: Chemical escalation in World War I. Bulletin of the Atomic Scientists, 27 (1):34–38, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:1974:CAM

[Coo74] C. Sharp Cook. A comment on “Los Angeles and Its Mistress Machine”. Bulletin of the Atomic Scientists, 30(2):23–24, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar73].

Cook:1975:EWH

[Coo75a] C. Sharp Cook. Editorial: What happens when the oil is gone? Bulletin of the Atomic Scientists, 31(6):7–9, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coote:1975:LDT

[Coo75b] G. E. Coote. Letters: The day they blew up …. Bulletin of the Atomic Scientists, 31(6):3, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch74b].

Cook:1976:DSW

[Coo76a] C. Sharp Cook. Don’t say we weren’t warned. Bulletin of the Atomic Scientists, 32(7):21–24, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:1976:LPP

[Coo76b] Charles Cook. Letter: A peace project. Bulletin of the Atomic Scientists, 32(10):5, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooke:1976:LRE

[Coo76c] H. R. Cooke, Jr. Letter: The role of experts. Bulletin of the Atomic Scientists, 32(8):2, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol76a].

Cooke:1977:LMP

[Coo77] H. R. Cooke, Jr. Letter: Man’s photo-finish. Bulletin of the Atomic Scientists, 33(1):3–4, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:1978:VIC

[Coo78] C. Sharp Cook. A view of Iceland: the country has all it can do to maximize its own energy resources. Bulletin of the Atomic Scientists, 34(3):34–36, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooke:1979:CDG

[Coo79] Dick Cooke. Commentary: The dating game. Bulletin of the Atomic Scientists, 35(9):65, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Koe79].

Cooper:1981:BRB

[Coo81] Chester L. Cooper. Book review: The Destruction of Nuclear Energy Facilities in War, by Bennett Ramberg. Bulletin of the Atomic Scientists, 37(5):45–46, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooper:1985:VIP

[Coo85] Chester L. Cooper. Vienna institute perseveres. Bulletin of the Atomic Scientists, 41(11):49–51, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooper:1990:SMF

[Coo90] Julian Cooper. Soviet military has a finger in every pie. Bulletin of the Atomic Scientists, 46(10):22–25, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cook:2004:RO

[Coo04] Dana Cook. Remembering Oppenheimer. Bulletin of the Atomic Scientists, 60(2):37–41, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooke:2006:RJW

[Coo06] Stephanie Cooke. Reprocessing: Just within reach? Bulletin of the Atomic Scientists, 62(4):14–17, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/4/14.

Cooper:2011:IFU

[Coo11] Mark Cooper. The implications of Fukushima: The US perspective. Bulletin of the Atomic Scientists, 67(4):8–13, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/4/8.

Cooper:2012:NSA

[Coo12] Mark Cooper. Nuclear safety and affordable reactors: Can we have both? Bulletin of the Atomic Scientists, 68(4):61–72, July/ August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/68/
4/61.

Coppock:1965:BRV

[Cop65] John O. Coppock. Book review: Vincent P. Rock, A Strategy of Interdependence. Bulletin of the Atomic Scientists, 21(2):40, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Copeland:1980:CWP

[Cop80] E. H. Copeland. Commentary: A welcome proposal. Bulletin of the Atomic Scientists, 36(8):62, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AP80].

Coryell:1953:CWN

[Cor53] Charles D. Coryell. Chaim Weizmann: Nov. 27, 1874–Nov. 9, 1952: The idealist, scientist, leader, and statesman. Bulletin of the Atomic Scientists, 9(1):31–32, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coryell:1956:PP

[Cor56] Charles C. Coryell. A project for peace. Bulletin of the Atomic Scientists, 12(6):201, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cornog:1958:CAH

[Cor58] William H. Cornog. The crisis in American high schools. Bulletin of the Atomic Scientists, 14(9):354–358, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Corning:1969:RRA

[Cor69] Erastus Corning II. Reports: The race for (automobile) space. Bulletin of the Atomic Scientists, 25(10):15–16, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Corradini:1978:UNC

[Cor78] Alessandro Corradini. United Nations convenes disarmament conference: Disarmament progress requires a new political climate. Bulletin of the Atomic Scientists, 34(4):14–17, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cortright:1984:CST

[Cor84] David Cortright. Commentary: SANE today. Bulletin of the Atomic Scientists, 40(9):54, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cortright:1992:MM

[Cor92] David Cortright. From the movement to the moment. Bulletin of the Atomic Scientists, 48(4):13–15, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cortright:1994:BRP

[Cor94] David Cortright. Book review: The power of public pressure: Peace Politics, by Paul Joseph. Bulletin of the Atomic Scientists, 50(2):56–58, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cortright:1995:BAL

[Cor95] David Cortright. Bulletin: Los Alamos’s little war with peace. Bulletin of the Atomic Scientists, 51(2):5–6, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cortright:1996:CI

[Cor96] David Cortright. The coming of incrementalism. Bulletin of the Atomic Scientists, 52(2):32–36, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi96].

Cosic:1965:TIC

[Ćos65] Dobrica Ćosić. Toward an intellectual community. Bulletin of the Atomic Scientists, 21(6):2–6, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translated by Jan Dekker.

Costner:2002:AD

[Cos02] Brian Costner. Access denied. Bulletin of the Atomic Scientists, 58(2):58–62, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/2/58.

Cotton:1952:SFE

[Cot52] Eugenie Cotton. Some French experiences. Bulletin of the Atomic Scientists, 8(7):242, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Cotter:1984:CCD

[Cot84] Donald R. Cotter. Commentary: Conventional defense in Europe. Bulletin of the Atomic Scientists, 40(4):48, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cote:2019:INA

[Cot19] Owen R. Cote Jr. Invisible nuclear-armed submarines, or transparent oceans? Are ballistic missile submarines still the best deterrent for the United States? Bulletin of the Atomic Scientists, 75(1):30–35, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1555998.

Cote:2022:OII

[Cot22] Owen R. Cote Jr. One if by invasion, two if by coercion: US military capacity to protect Taiwan from China. Bulletin of the Atomic Scientists, 78(2):65–72, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2038882.

CBASA:1948:BAS

[Cou48] Council of the British Atomic Scientists Association. British atomic scientists’s statement on international control of atomic energy. Bulletin of the Atomic Scientists, 4(8):236, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scientists:1950:FSS

[Cou50] Council of the Federation of American Scientists. FAS statement on Science Foundation Bill. Bulletin of the Atomic Scientists, 6 (6):190, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coulomb:1952:SFE

[Cou52a] J. Coulomb. Some French experiences. Bulletin of the Atomic Scientists, 8(7):241, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

CFAS:1952:UTA

[Cou52b] Council of the Federation of American Scientists. The use of tactical atomic weapons. Bulletin of the Atomic Scientists, 8(5): 137, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scientists:1954:PAN

[Cou54] Council of the Federation of American Scientists. Prospective appointments to the National Science Board. Bulletin of the Atomic Scientists, 10(3):104, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

CFAS:1957:FPS

[Cou57] Council of the Federation of American Scientists. FAS policy statement on arms limitation, February 3, 1957. Bulletin of the Atomic Scientists, 13(4):138, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coulomb:1971:GUS

[Cou71] Jean Coulomb. Good use of scientists. Bulletin of the Atomic Scientists, 27(1):39–41, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coulam:1977:ISW

[Cou77] Robert F. Coulam. Inter-service weapons rivalry. Bulletin of the Atomic Scientists, 33(6):25–36, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Coulter:1981:CMB

[Cou81] N. A. Coulter, Jr. Commentary: MAD better than NUT. Bulletin of the Atomic Scientists, 37(9):64, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Buc81].

Cowen:1965:BBR

[Cow65] Robert C. Cowen. Books: The Rise and Fall of the Space Age, by Edwin Diamond; The Moon-Doggle: Domestic and International Implications of the Space Race, by Amitai Etzioni. Bulletin of the Atomic Scientists, 21(4):38, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cowlan:1971:CGT

[Cow71] Bert Cowlan. Communications: Gofman, Tamplin, and the AEC. Bulletin of the Atomic Scientists, 27(2):4, February 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TG70].

Cox:1986:LGP

[Cox86] Eric Cox. Letter: Global peace-keeping institutions. Bulletin of the Atomic Scientists, 42(1):53–54, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro85g].

Crutchfield:1962:CF

[CP62] James Crutchfield and Giulio Pontecorvo. Crisis in the fisheries. Bulletin of the Atomic Scientists, 18(9):18–20, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cahn:1993:TBT

[CP93a] Anne Hessing Cahn and John Prados. Team-B: the trillion dollar experiment. Bulletin of the Atomic Scientists, 49(3):22–31, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-13610752.

html. See corrections [Ano93-31].

Christiansen:1993:SUC

[CP93b] Drew Christiansen and Gerard F. Powers. Sanctions: Unintended consequences. Bulletin of the Atomic Scientists, 49(9):41–45, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Carter:1998:GYM

[CP98] Luther J. Carter and Thompson H. Pigford. Getting Yucca Mountain right. Bulletin of the Atomic Scientists, 54(2):56–61, March/ April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Collina:1999:TBT

[CP99] Tom Z. Collina and Christopher Paine. Test ban treaty: Let’s finish the job. Bulletin of the Atomic Scientists, 55(4):24–27, July/ August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
4/24.

Cameron:2001:PFH

[CPV01] Gavin Cameron, Jason Pate, and Kathleen Vogel. Planting fear: How real is the threat of agricultural terrorism? Bulletin of the Atomic Scientists, 57(5):38–44, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/38.

Croke:1971:ARM

[CR71] E. J. Croke and J. J. Roberts. Air resource management and regional planning. Bulletin of the Atomic Scientists, 27(2):8–12, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Callen:1981:MW

[CR81] Earl Callen and Anthony Ralston. To Moscow for the weekend. Bulletin of the Atomic Scientists, 37(9):38–42, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Eva82].

Cotta-Ramusino:1999:UQ

[CR99] Paolo Cotta-Ramusino. The unasked question. Bulletin of the Atomic Scientists, 55(4):42–45, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/4/42.

Carrasco:2017:CES

[CR17] Camila Carrasco and Robert Rosner. The Chilean electricity sector confronts climate change. Bulletin of the Atomic Scientists, 73(6):395–403, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1388677.

Crane:1949:LE

[Cra49] M. B. Crane. Lysenko’s experiments. Bulletin of the Atomic Scientists, 5(5):147–149, 156, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crary:1960:RA

[Cra60] A. P. Crary. A report on Antarctica. Bulletin of the Atomic Scientists, 16(10):376–381, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crawford:1961:LET

[Cra61] Curtis Crawford. Letter to the Editor: Toward a more creative diplomacy. Bulletin of the Atomic Scientists, 17(1):0, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crary:1964:AIC

[Cra64] A. P. Crary. Antarctica: International cooperation: The international laboratory. Bulletin of the Atomic Scientists, 20(1):27–30, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cranberg:1967:PMB

[Cra67a] Lawrence Cranberg. The P-M bomb. Bulletin of the Atomic Scientists, 23(10):37–40, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See sequel [Cra68].

Crawford:1967:CWM

[Cra67b] Eric W. Crawford. Comments: Withdraw missile submarines? Bulletin of the Atomic Scientists, 23(9):39, November 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cranberg:1968:ES

[Cra68] Lawrence Cranberg. Ethics of scientists. Bulletin of the Atomic Scientists, 24(7):39–40, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cra67a].

Craddock:1970:PSL

[Cra70a] Campbell Craddock. Part 2: Science laboratory: Antarctic geology and Gondwanaland. Bulletin of the Atomic Scientists, 26 (10):33–39, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Crary:1970:LLA

[Cra70b] A. P. Crary. The long look ahead. Bulletin of the Atomic Scientists, 26(10):100–104, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Craig:1972:CM

[Cra72] Eugene A. Craig. Communications: Mariners. Bulletin of the Atomic Scientists, 28(10):3, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cranston:1976:LSC

[Cra76] Frederick P. Cranston. Letter: Solar crash program. Bulletin of the Atomic Scientists, 32(7):5, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1982:CSU

[Cra82a] James Cracraft. Commentary from the Soviet Union: From the Russian past to the Soviet present. Bulletin of the Atomic Scientists, 38(1):8–12, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1982:ENS

[Cra82b] James Cracraft. [Editorial note on the Soviet economy and the consumer]. Bulletin of the Atomic Scientists, 38(4):24, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1982:SEN

[Cra82c] James Cracraft. [Special Editor’s note on “The Soviet Union: The power elite”]. Bulletin of the Atomic Scientists, 38(8):24, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracroft:1982:NSE

[Cra82d] James Cracroft. [Note from the special editor on Soviet post-war architecture and planning]. Bulletin of the Atomic Scientists, 38 (3):14, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1983:BNB

[Cra83a] James Cracraft. Book note: Soviet Armed Forces Review Annual, edited by David R. Jones. Bulletin of the Atomic Scientists, 39 (7):51, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1983:BRB

[Cra83b] James Cracraft. Book reviews: The Soviet Art of War: Doctrine, Strategy, and Tactics, by Harriet Fast Scott and William F. Scott; Soviet Perceptions of Military Power: The Interaction of Theory and Practice, by John J. Dziak; Soviet Marxism and Nuclear War: An International Debate, by John Somerville and others; Diplomacy of Power: Soviet Armed Forces as a Political Instrument, by Stephen S. Kaplan and others; The Soviet Estimate: U.S. Intelligence Analysis and Russian Military Strength, by John Prados. Bulletin of the Atomic Scientists, 39(1):34–36, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1983:SUT

[Cra83c] James Cracraft, editor. The Soviet Union today. Educational Foundation for Nuclear Science, Chicago, IL, USA, 1983. ISBN 0-941682-06-4 (paperback). ix + 348 pp. LCCN DK4 .S56 1983.

Cracraft:1983:SEN

[Cra83d] James Cracraft. [Special Editor’s note on “The Soviet Union: The conscripts”]. Bulletin of the Atomic Scientists, 39(5):28, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:AUS

[Cra84a] James Cracraft. Agenda 1984: U.S.–Soviet relations. Bulletin of the Atomic Scientists, 40(1):3–5, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:BNB

[Cra84b] James Cracraft. Book note: Soviet Power and Western Negotiating Policies, edited by Uwe Nerlich; Military Power in the Competition over Europe; Vol. 2, The Western Panacea: Constraining Soviet Power Through Negotiation. Bulletin of the Atomic Scientists, 40(6):39, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:BRBc

[Cra84c] James Cracraft. Book review: The Making of America’s Soviet Policy, edited by Joseph S. Nye, Jr. Bulletin of the Atomic Scientists, 40(9):49–50, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:BRBa

[Cra84d] James Cracraft. Book review: The Soviet Union and the Arms Race, by David Holloway. Bulletin of the Atomic Scientists, 40(3): 40–43, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:BRBb

[Cra84e] James Cracraft. Book reviews: The Russians and Reagan, by Strobe Talbott; The Dilemma of Reform in the Soviet Union, by Timothy J. Colton. Bulletin of the Atomic Scientists, 40(8):40– 42, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:MTW

[Cra84f] James Cracraft. Militarization of the Third World. Bulletin of the Atomic Scientists, 40(5):31–32, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1984:PWR

[Cra84g] James Cracraft. Perspective: Where Russians and Americans meet. Bulletin of the Atomic Scientists, 40(4):5–6, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1985:OUS

[Cra85] James Cracraft. 1985 outlook: U.S.–Soviet relations. Bulletin of the Atomic Scientists, 41(1):7–8, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1986:STP

[Cra86a] James Cracraft. A Soviet turning point. Bulletin of the Atomic Scientists, 42(2):8–12, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [EG86].

Cracraft:1986:UTC

[Cra86b] James Cracraft. Unofficial thaw in the Cold War. Bulletin of the Atomic Scientists, 42(5):20–22, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1987:GRA

[Cra87] James Cracraft. The Gorbachev regime after two years. Bulletin of the Atomic Scientists, 43(4):31–33, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cracraft:1988:BRB

[Cra88] James Cracraft. Book review: Breakthrough: Emerging New Thinking, edited by Anatoly Gromyko and Martin Hellman. Bulletin of the Atomic Scientists, 44(6):50–51, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cranston:1996:GOE

[Cra96] Alan Cranston. Guest opinion: Even the generals agree. Bulletin of the Atomic Scientists, 52(6):4, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Craig:1998:LMP

[Cra98] Paul Craig. Letter: MOX plans all wrong. Bulletin of the Atomic Scientists, 54(4):65–66, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cranston:2001:CA

[Cra01] Alan Cranston. Commit to abolition. Bulletin of the Atomic Scientists, 57(1):39–41, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/1/39.

Chesser:2009:PTI

[CRB^{+}09] Ronald K. Chesser, Brenda E. Rodgers, Mikhail Bondarkov, Esmail Shubber, and Carleton J. Phillips. Piecing together Iraq’s nuclear legacy. Bulletin of the Atomic Scientists, 65(3):19–33, May 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/3/19.

Crewe:1963:ICR

[Cre63] Albert V. Crewe. International cooperation: A red carpet. Bulletin of the Atomic Scientists, 19(7):25–26, September 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crewe:1968:BBS

[Cre68] Albert V. Crewe. Books: Crime and Science, by Jürgen Thorwald. Bulletin of the Atomic Scientists, 24(5):37–38, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Creutz:1970:NPR

[Cre70] Edward Creutz. Nuclear power: Rise of an industry. Bulletin of the Atomic Scientists, 26(6):75–82, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crewe:1974:CEC

[Cre74] Albert V. Crewe. Communications: Energy crisis and basic research. Bulletin of the Atomic Scientists, 30(2):2, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cremins:1990:OSM

[Cre90] Thomas E. Cremins. Open season on Moscow. Bulletin of the Atomic Scientists, 46(4):5, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crigger:1994:YTK

[Cri94] Bette-Jane Crigger. Yes, they knew better. Bulletin of the Atomic Scientists, 50(2):29, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crow:1958:RMB

[Cro58] James F. Crow. Radiation and man: Basic data: Genetic effects of radiation. Bulletin of the Atomic Scientists, 14(1):19–22, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crosby:1964:NDE

[Cro64] H. Ashton Crosby. Nonnuclear defense of Europe. Bulletin of the Atomic Scientists, 20(4):30–31, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crocker:1965:PHC

[Cro65a] Thomas D. Crocker. In Polk & Hillsborough Counties, Florida. Bulletin of the Atomic Scientists, 21(6):17–19, June 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crosby:1965:BBW

[Cro65b] H. A. Crosby. Books: Conventional Warfare in the Nuclear Age, by Otto Heilbrunn; Limited War and American Defense Policy, by Seymour Deitchman. Bulletin of the Atomic Scientists, 21 (7):34–36, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crosby:1965:NT

[Cro65c] H. Ashton Crosby. Nato tomorrow. Bulletin of the Atomic Scientists, 21(5):18–21, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crow:1967:BBN

[Cro67] James P. Crow. Books: Heredity and the Nature of Man, by Theodosius Dobzhansky; Internal Factors in Evolution, by Lancelot L. Whyte. Bulletin of the Atomic Scientists, 23(2):22–23, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crosby:1968:CCS

[Cro68] H. Ashton Crosby. Current comments: On Schelling and the Selective Service. Bulletin of the Atomic Scientists, 24(2):27–28, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ACC^{+}67, Sch68a].

Cronin:1986:PCS

[Cro86] James W. Cronin. Perspective: The case for the super collider. Bulletin of the Atomic Scientists, 42(5):8–11, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pri86a, Vog87].

Cronin:2007:RS

[Cro07] James W. Cronin. Relatively speaking. Bulletin of the Atomic Scientists, 63(5):55–56, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/5/55.

Crowder:2016:ENa

[Cro16a] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 72(3):185, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crowder:2016:ENb

[Cro16b] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 72(4):259–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crowder:2016:ENc

[Cro16c] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 72(5):339–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crowder:2016:ENd

[Cro16d] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 72(6):359, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Crowder:2017:ENa

[Cro17a] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 73(1):58, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2016.1264214.

Crowder:2017:ENb

[Cro17b] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 73(2):127, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1288449.

Crowder:2017:EN

[Cro17c] Lucien Crowder. Editor’s note. Bulletin of the Atomic Scientists, 73(3):196, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com//
doi/abs/10.1080/00963402.2017.1315081.

Crowder:2017:NCT

[Cro17d] Lucien Crowder. Noam Chomsky: In the Trump era, severe threats to “organized human life”. Bulletin of the Atomic Scientists, 73(6):358–363, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1388659.

Crowder:2018:VBI

[Cro18] Lucien Crowder. Ventilator blues: Infectious disease expert Tom Inglesby on the next major pandemic. Bulletin of the Atomic Scientists, 74(5):340–347, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2018.1507794.

Crutchfield:1967:ECN

[Cru67] James A. Crutchfield. Economic considerations [of Nawapa]. Bulletin of the Atomic Scientists, 23(7):17–21, September 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cohen:1963:LEP

[CS63] John Cohen and J. David Singer. Letter to the Editor: Peace research. Bulletin of the Atomic Scientists, 19(5):35–36, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sin63c].

Callen:1975:ASE

[CS75] Earl Callen and Edward A. Stein. Abuses of scientific exchanges. Bulletin of the Atomic Scientists, 31(2):32–35, February 1975.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Charap:2016:URR

[CS16] Samuel Charap and Jeremy Shapiro. US–Russian relations: The middle cannot hold. Bulletin of the Atomic Scientists, 72(3):150– 155, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cooper:2017:TCA

[CS17] Zack Cooper and Andrew Shearer. Thinking clearly about China’s layered Indo–Pacific strategy. Bulletin of the Atomic Scientists, 73(5):305–311, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1364005.

Carter:1995:LUT

[CSF95] Luther J. Carter and Kristin Shrader-Frechette. Letters: Use test site as waste site. Bulletin of the Atomic Scientists, 51(1): 65–67, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SF94].

Chase:1971:CMT

[CSG71] Stuart Chase, Homer W. Smith, and Jay B. Gaskill. Communications: Mounting tide of unreason. Bulletin of the Atomic Scientists, 27(9):3, 32, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab71c].

Callaham:1978:CBW

[CT78] Michael B. Callaham and Kosta Tsipis. Commentary: Biological warfare and recombinant DNA. Bulletin of the Atomic Scientists, 34(9):11, 50, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Culver:1979:DI

[Cul79] John C. Culver. Dangerous illusions. Bulletin of the Atomic Scientists, 35(7):39–40, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Culp:2008:DL

[Cul08] David Culp. The Domenici Legacy. Bulletin of the Atomic Scientists, 64(5):22–25, November/December 2008. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/5/22.

Cullenward:2014:HCC

[Cul14] Danny Cullenward. How California’s carbon market actually works. Bulletin of the Atomic Scientists, 70(5):35–44, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cumings:1992:MB

[Cum92a] Bruce Cumings. MacArthur’s bombs. Bulletin of the Atomic Scientists, 48(3):21, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cumings:1992:STK

[Cum92b] Bruce Cumings. Spring thaw for Korea’s Cold War? Bulletin of the Atomic Scientists, 48(3):14–23, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cummings:1994:GOM

[Cum94] Bruce Cummings. Guest opinion: A monstrous idea. Bulletin of the Atomic Scientists, 50(5):4, September/October 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cumings:1997:MB

[Cum97a] Bruce Cumings. The military balance. Bulletin of the Atomic Scientists, 53(1):44, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cumings:1997:WCW

[Cum97b] Bruce Cumings. Where the Cold War never ends. Bulletin of the Atomic Scientists, 53(1):40–46, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cummings:1998:BRN

[Cum98] Bruce Cummings. Book review: The near-war: The Two Koreas: A Contemporary History, by Don Oberdorfer. Bulletin of the Atomic Scientists, 54(3):62–64, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cumings:1999:CTL

[Cum99a] Bruce Cumings. China through the looking glass. Bulletin of the Atomic Scientists, 55(5):30–37, September/October 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/5/30.

Cumings:1999:DCI

[Cum99b] Bruce Cumings. Don’t count on it. Bulletin of the Atomic Scientists, 55(3):72–74, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/3/72.

Cumings:2001:LHB

[Cum01a] Bruce Cumings. A little history is a bad thing. Bulletin of the Atomic Scientists, 57(2):70–71, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/70.

Cumings:2001:SCA

[Cum01b] Bruce Cumings. Stay the course on Asia. Bulletin of the Atomic Scientists, 57(1):29–30, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/29.

Cumings:2002:HCH

[Cum02] Bruce Cumings. Hegemons can’t help it. Bulletin of the Atomic Scientists, 58(3):67–68, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/3/67.

Cumings:2003:NL

[Cum03] Bruce Cumings. The next in line. Bulletin of the Atomic Scientists, 59(1):24–25, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/24.

Cumings:2004:MB

[Cum04] Bruce Cumings. Mao and beyond. Bulletin of the Atomic Scientists, 60(4):67–69, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/4/67.

Cumings:2006:NE

[Cum06] Bruce Cumings. Northern exposure. Bulletin of the Atomic Scientists, 62(2):63, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/2/63.

Cumings:2015:GNK

[Cum15] Bruce Cumings. Getting North Korea wrong. Bulletin of the Atomic Scientists, 71(4):64–76, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cunliffe:1952:BRM

[Cun52] Marcus Cunliffe. The British reactions to the McCarran Acts. Bulletin of the Atomic Scientists, 8(7):233, 256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Cunningham:1988:LTH

[Cun88] Clarence M. Cunningham. Letter: Teller’s heavy hand. Bulletin of the Atomic Scientists, 44(8):52, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu88].

Cushing:1948:SAA

[Cus48] Richard J. Cushing. A spiritual approach to the atomic age. Bulletin of the Atomic Scientists, 4(7):222–224, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cushman:1949:FVS

[Cus49] Robert E. Cushman. Freedom vs. security. Bulletin of the Atomic Scientists, 5(3):69–72, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cutler:1958:LEE

[Cut58] Phil Cutler. Letter to the Editor: An exchange of people. Bulletin of the Atomic Scientists, 14(4):156, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cutler:1975:LTH

[Cut75] Hymie Cutler. Letters: Too high a price. Bulletin of the Atomic Scientists, 31(2):4, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pri74, Pri75a].

Calder:1981:CVP

[CW81] Nigel Calder and Andrew Wilson. Commentary from Vienna: Pugwash meets with the press. Bulletin of the Atomic Scientists, 37(3):58, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Cerf:2023:IYW

[CW23] Moran Cerf and Adam Waytz. If you worry about humanity, you should be more scared of humans than of AI. Bulletin of the Atomic Scientists, 79(5):289–292, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2245242.

Coates:1981:CEN

[CY81] Ken Coates and Elizabeth Young. Commentary: European Nuclear Disarmament Campaign. Bulletin of the Atomic Scientists, 37(9):57–60, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [You81].

Carr:2014:SUE

[CYP14] Wylie Carr, Laurie Yung, and Christopher Preston. Swimming upstream: Engaging the American public early on climate engineering. Bulletin of the Atomic Scientists, 70(3):38–48, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Campbell:1995:LRM

[CZ95] Alex Campbell and Tom Zamora Collina. Less regulation, more enterprise. Bulletin of the Atomic Scientists, 51(3):44–45, May/ June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DAmato:1970:EDL

[D’A70] Anthony A. D’Amato. Environmental degradation and legal action. Bulletin of the Atomic Scientists, 26(3):24–26, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daalder:1987:TDI

[Daa87] Ivo H. Daalder. A tactical defense initiative for Western Europe? Bulletin of the Atomic Scientists, 43(4):34–39, May 1, 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daddario:1967:CFS

[Dad67] Emilio Q. Daddario. Congress faces space policies. Bulletin of the Atomic Scientists, 23(5):11–16, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daddario:1971:NSP

[Dad71] Emilio Q. Daddario. National science policy — prelude to global cooperation. Bulletin of the Atomic Scientists, 27(6):21–24, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dafler:1964:LEE

[Daf64] James R. Dafler. Letter to the Editor: Ethics. Bulletin of the Atomic Scientists, 20(10):22, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hay64].

Dahlberg:1977:CPR

[Dah77] Richard C. Dahlberg. The case for plutonium recycle: Let’s not immobilize ourselves in trying to solve problems that may not be as formidable as we think. Bulletin of the Atomic Scientists, 33 (5):58–59, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dahlberg:1978:WPC

[Dah78] Richard C. Dahlberg. Weapons proliferation and criteria for evaluating nuclear fuel cycles. Bulletin of the Atomic Scientists, 34 (1):38–42, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lit78].

Daleus:1975:MNO

[Dal75] Lennart Daléus. A moratorium in name only. Bulletin of the Atomic Scientists, 31(8):27–33, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daly:1982:BRB

[Dal82] Herman E. Daly. Book review: The Ultimate Resource, by Julian Simon. Bulletin of the Atomic Scientists, 38(1):39–42, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SD83].

Dallin:1983:SUM

[Dal83] Alexander Dallin. The Soviet Union: The making of foreign policy. Bulletin of the Atomic Scientists, 39(7):27–31, August/ September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daly:1986:PTN

[Dal86] Herman E. Daly. Population: Toward a new economic model. Bulletin of the Atomic Scientists, 42(4):42–44, April 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daley:1992:CUS

[Dal92a] Tad Daley. Can the U.N. stretch to fit its future? Bulletin of the Atomic Scientists, 48(3):38–42, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daley:1992:CSS

[Dal92b] Tad Daley. “Continuing” the Soviet seat. Bulletin of the Atomic Scientists, 48(3):41, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daley:2003:WSM

[Dal03] Tad Daley. When speed matters. Bulletin of the Atomic Scientists, 59(3):22–25, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/3/22.

Dalton:2017:HIH

[Dal17] Melissa G. Dalton. How Iran’s hybrid-war tactics help and hurt it. Bulletin of the Atomic Scientists, 73(5):312–315, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1362904.

Daly:2022:CUS

[Dal22] Robert Daly. China and the United States: It’s a Cold War, but don’t panic. Bulletin of the Atomic Scientists, 78(2):59–64, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2038881.

Dam:1961:BSC

[Dam61] Kenneth W. Dam. Books: Scientists and conflict of interest: Conflict of Interest and Federal Service. Bulletin of the Atomic Scientists, 17(8):342, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dam:1962:BBN

[Dam62] Kenneth W. Dam. Books: Symposium on Next Steps in Extending the Rule of Law. Bulletin of the Atomic Scientists, 18(5):38, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniels:1946:APP

[Dan46] Farrington Daniels. Atomic power production. Bulletin of the Atomic Scientists, 1(7):13–14, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniels:1948:ANL

[Dan48] Farrington Daniels. The Argonne National Laboratory. Bulletin of the Atomic Scientists, 4(6):177–180, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dancoff:1952:DNR

[Dan52] S. M. Dancoff. Does the neutrino Really exist? Bulletin of the Atomic Scientists, 8(5):138–141, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniels:1954:IGS

[Dan54a] Farrington Daniels. [Interference with government service and personal justice]. Bulletin of the Atomic Scientists, 10(3):71, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniels:1954:TAP

[Dan54b] Farrington Daniels. Technical assistance programs — Part II: Implications of Point Four for the scientists. Bulletin of the Atomic Scientists, 10(4):125, 144, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniels:1957:BBQ

[Dan57] Farrington Daniels. Books: Atomic Quest, by Arthur Compton. Bulletin of the Atomic Scientists, 13(4):143–145, April 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniels:1970:PPL

[Dan70] Paul C. Daniels. Part 1: Political laboratory: The Antarctic Treaty. Bulletin of the Atomic Scientists, 26(10):11–15, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Danielli:1972:CMC

[Dan72] James F. Danielli. Can man control his biological evolution? — a symposium on genetic engineering: IV. Artificial synthesis of new life forms. Bulletin of the Atomic Scientists, 28(10):20–24, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Danielsson:1990:CHD

[Dan90a] Bengt Danielsson. Cousteau out of his depth? Bulletin of the Atomic Scientists, 46(2):25, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Danielsson:1990:PPL

[Dan90b] Bengt Danielsson. Poisoned Pacific: the legacy of French nuclear testing. Bulletin of the Atomic Scientists, 46(2):22–31, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Danielsson:1990:TS

[Dan90c] Bengt Danielsson. Tahiti secrets. Bulletin of the Atomic Scientists, 46(2):27, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Danielsson:1991:RNT

[Dan91] Bengt Danielsson. Report: Nuclear testing: French slam “open door” on Greenpeace. Bulletin of the Atomic Scientists, 47(2): 6–7, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dando:2005:BC

[Dan05] Malcolm Dando. The bioterrorist cookbook. Bulletin of the Atomic Scientists, 61(6):34–39, November/December 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/6/34.

Dando:2016:ACC

[Dan16] Malcolm Dando. Arms control in the 21st century: A holistic approach. Bulletin of the Atomic Scientists, 72(3):194–197, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dando:2020:BC

[Dan20] Malcolm Dando. 2005: The bioterrorist cookbook. Bulletin of the Atomic Scientists, 76(6):423–427, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847522.

Darlington:1951:CSS

[Dar51] C. D. Darlington. The conflict of science and society. Bulletin of the Atomic Scientists, 7(1):9–12, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Darlington:1957:FRA

[Dar57] C. D. Darlington. Freedom and responsibility in academic life. Bulletin of the Atomic Scientists, 13(4):131–134, April 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Darwin:1958:PP

[Dar58] Charles Darwin. Population problems. Bulletin of the Atomic Scientists, 14(8):322–324, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Darmstadter:1971:GDM

[Dar71] Joel Darmstadter. GNP does matter. Bulletin of the Atomic Scientists, 27(10):46–48, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dardi:1986:LOS

[Dar86] Peter S. Dardi. Letter: Origins of SDI. Bulletin of the Atomic Scientists, 42(2):47–48, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Darchiyev:1991:BRW

[Dar91] Alexander N. Darchiyev. Book review: Where left is right: Farewell Perestroika: A Soviet Chronicle, by Boris Kagarlitsky. Bulletin of the Atomic Scientists, 47(2):41–42, March 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dash:1966:WRL

[Das66] J. Gregory Dash. Where responsibility lies. Bulletin of the Atomic Scientists, 22(1):35–37, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dash:1968:NFP

[Das68] J. Gregory Dash. Nuclear-free and people-free zones. Bulletin of the Atomic Scientists, 24(7):34–35, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dash:1985:LWL

[Das85] J. G. Dash. Letter: Weapons laboratories. Bulletin of the Atomic Scientists, 41(2):54, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DeW84] and comment [McC85].

Dasgupta:1996:BRO

[Das96] S. Dasgupta. Book review: Operation Brasstacks: Brasstacks and beyond: Perception and Management of Crisis in South Asia, by Kanti P. Bajpai, P. R. Chari, Pervaiz Iqbal Cheema, Stephen P. Cohen, and Sumit Ganguly. Bulletin of the Atomic Scientists, 52(1):57–58, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davies:1948:SRC

[Dav48] T. H. Davies. “Security risk” cases — a vexed question. Bulletin of the Atomic Scientists, 4(7):193–194, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davies:1951:SAE

[Dav51] T. H. Davies. “Soviet atomic espionage”. Bulletin of the Atomic Scientists, 7(5):143–148, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davies:1952:SBE

[Dav52] V. R. E. Davies. Some British experiences. Bulletin of the Atomic Scientists, 8(7):231, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Davis:1953:CS

[Dav53] Watson Davis. Communicating science. Bulletin of the Atomic Scientists, 9(9):337–340, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidson:1955:LEB

[Dav55] William C. Davidson. Letter to the Editor: [Bulletin clock should be set back]. Bulletin of the Atomic Scientists, 11(7):271, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davenport:1957:LEC

[Dav57a] John H. Davenport. Letter to the Editor: A Christian answer to war. Bulletin of the Atomic Scientists, 13(4):152–153, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [McC56, Koc57].

Davidon:1957:MRC

[Dav57b] William C. Davidon. The modern Roman circus. Bulletin of the Atomic Scientists, 13(9):331–332, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidon:1957:SSU

[Dav57c] William C. Davidon. Soviet satellites — U.S. reactions. Bulletin of the Atomic Scientists, 13(10):357–358, December 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidon:1958:CE

[Dav58] William C. Davidon. Contrasts in education. Bulletin of the Atomic Scientists, 14(4):154–156, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidon:1960:NWT

[Dav60] William Davidon. Nuclear weapons test ban negotiations. Bulletin of the Atomic Scientists, 16(8):334–335, October 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:1965:LEB

[Dav65] Joyce P. Davis. Letter to the Editor: Bodega Head. Bulletin of the Atomic Scientists, 21(8):27–28, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hed65, HH65].

Davis:1968:LO

[Dav68] Nuel Pharr Davis. Lawrence and Oppenheimer. Simon and Schuster, New York, NY, USA, 1968. 384 pp. LCCN QC16.L36 D38.

Davis:1969:BBM

[Dav69] Morton Davis. Books: Mathematical Modes of Arms Control & Disarmament, by Thomas L. Saaty. Bulletin of the Atomic Scientists, 25(8):42, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davenport:1972:CEC

[Dav72] John H. Davenport. Communications: On the energy crisis. Bulletin of the Atomic Scientists, 28(3):2, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spi71a, Spi72b].

Davenport:1974:LLV

[Dav74] John H. Davenport. Letter: Lost vision. Bulletin of the Atomic Scientists, 30(8):3, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ols74a].

Davis:1976:LSA

[Dav76a] Bernard D. Davis. Letter: Shallow analysis. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Davis:1976:UFH

[Dav76b] Saville Davis. UNESCO: a future hanging in the balance. Bulletin of the Atomic Scientists, 32(4):16–21, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidon:1977:MOP

[Dav77a] Ann Morrissett Davidon. Macho obstacles to peace: how violent conflict is nurtured by human misconceptions about the nature of power and weakness. Bulletin of the Atomic Scientists, 33(6): 22–24, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Tob77, Ago77, Mau77].

Davison:1977:LES

[Dav77b] J. Davison. Letter: Enough shallow talk. Bulletin of the Atomic Scientists, 33(7):6, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davison:1977:LOH

[Dav77c] Jerome Davison. Letter: The only hope. Bulletin of the Atomic Scientists, 33(5):6, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidon:1979:CUA

[Dav79a] Ann Morrissett Davidon. Commentary: The U.S. anti-nuclear movement. Bulletin of the Atomic Scientists, 35(10):45–48, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davies:1979:BRB

[Dav79b] David Davies. Book reviews: Nuclear Explosions and Earthquakes: The Parted Veil, by Bruce Bolt; Monitoring Underground Nuclear Explosions, by Ola Dahlman and Hans Israelson. Bulletin of the Atomic Scientists, 35(2):40–44, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:1979:CIR

[Dav79c] P. R. Davis. Commentary: If reactors are targets. Bulletin of the Atomic Scientists, 35(1):62–63, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davison:1982:CAN

[Dav82] Jerome Davison. Commentary: Accidental nuclear war. Bulletin of the Atomic Scientists, 38(1):49, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davidon:1983:CIA

[Dav83] Ann Morrissett Davidon. Commentary: It all depends on definitions. Bulletin of the Atomic Scientists, 39(4):46–47, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HC83].

Davis:1986:PHB

[Dav86a] Kingsley Davis. Population: The history of birth and death. Bulletin of the Atomic Scientists, 42(4):20–23, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:1986:LO

[Dav86b] Nuel Pharr Davis. Lawrence and Oppenheimer. The Da Capo series in science; A Da Capo paperback. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1986. ISBN 0-306-80280-5 (paperback). 384 pp. LCCN QC16.L36 D38 1986. US$11.95.

Davydov:1992:NFB

[Dav92] Valeri Davydov. Nyet to full battle dress. Bulletin of the Atomic Scientists, 48(6):33–37, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:1993:SSA

[Dav93] Jennifer Davis. Sanctions: Squeezing apartheid. Bulletin of the Atomic Scientists, 49(9):16–19, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:1994:FMN

[Dav94] Mary Byrd Davis. The French mess nucléaire. Bulletin of the Atomic Scientists, 50(4):48–53, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

David:2005:CA

[Dav05] Leonard David. The clutter above. Bulletin of the Atomic Scientists, 61(4):32–37, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/4/32.

David:2007:ANG

[Dav07] Bosco David. All nations great and small. Bulletin of the Atomic Scientists, 63(2):53–54, March/April 2007. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/2/53.

Dawson:1968:IR

[Daw68] Laurence Dawson. Individual responsibility. Bulletin of the Atomic Scientists, 24(5):56, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1974:BAS

[Day74a] Samuel H. Day, Jr. Bulletin of the Atomic Scientists: Statement of purpose. Bulletin of the Atomic Scientists, 30(4):2–3, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1974:EAH

[Day74b] Samuel H. Day, Jr. Editorial: An appeal for help. Bulletin of the Atomic Scientists, 30(9):1, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pri75d] and response [Day75l].

Day:1974:ENC

[Day74c] Samuel H. Day, Jr. Editorial: Needed: A consumption-control bill. Bulletin of the Atomic Scientists, 30(6):37, June 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LSB74].

Day:1974:ESC

[Day74d] Samuel H. Day, Jr. Editorial: Scant cause for reassurance. Bulletin of the Atomic Scientists, 30(8):4, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Sch75b, Coh75a].

Day:1974:EWR

[Day74e] Samuel H. Day, Jr. Editorial: We re-set the clock. Bulletin of the Atomic Scientists, 30(7):4–5, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1974:MM

[Day74f] Samuel H. Day, Jr. Minutes to midnight. Bulletin of the Atomic Scientists, 30(6):8–9, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:NES

[Day75a] M. C. Day. Nuclear energy: a second round of questions. Bulletin of the Atomic Scientists, 31(10):52–59, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ner76].

Day:1975:BRC

[Day75b] Samuel H. Day, Jr. Book review: Case study of a meltdown: We Almost Lost Detroit, by John G. Fuller. Bulletin of the Atomic Scientists, 31(9):48, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:BBD

[Day75c] Samuel H. Day, Jr. Books: Advice and Dissent: Scientists in the Political Arena, by Joel Primack and Frank von Hippel. Bulletin of the Atomic Scientists, 31(2):47–48, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:EA

[Day75d] Samuel H. Day, Jr. Editorial: An agenda for 1975. Bulletin of the Atomic Scientists, 31(1):5–6, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:ESS

[Day75e] Samuel H. Day, Jr. Editorial: From Saigon to the Sinai. Bulletin of the Atomic Scientists, 31(8):5, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:ELP

[Day75f] Samuel H. Day, Jr. Editorial: Let’s put the troops to work. Bulletin of the Atomic Scientists, 31(4):4–5, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and reply [MD75, Squ75, Dum75].

Day:1975:EOU

[Day75g] Samuel H. Day, Jr. Editorial: Our unfinished business. Bulletin of the Atomic Scientists, 31(10):8–9, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ell76, Hal76a].

Day:1975:EPN

[Day75h] Samuel H. Day, Jr. Editorial: Prospects of nuclear violence: The drift becomes a rush. Bulletin of the Atomic Scientists, 31(7):6– 7, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:EBB

[Day75i] Samuel H. Day, Jr. Editorial: The bursting of the bubbles: No cause for lament. Bulletin of the Atomic Scientists, 31(6):6–7, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:EF

[Day75j] Samuel H. Day, Jr. Editor’s foreword. Bulletin of the Atomic Scientists, 31(7):15, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:LW

[Day75k] Samuel H. Day, Jr. The lessons of Watergate. Bulletin of the Atomic Scientists, 31(3):3–4, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fru75].

Day:1975:LRa

[Day75l] Samuel H. Day, Jr. Letters: Response. Bulletin of the Atomic Scientists, 31(1):2, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch74b, Mye75].

Day:1975:LRb

[Day75m] Samuel H. Day, Jr. Letters: Response. Bulletin of the Atomic Scientists, 31(1):3, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day74b, Pri75d].

Day:1975:OBI

[Day75n] Samuel H. Day, Jr. Ominous breakthrough [on isotope separation]. Bulletin of the Atomic Scientists, 31(6):29, June 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1975:SP

[Day75o] Samuel H. Day, Jr. The scientist in politics. Bulletin of the Atomic Scientists, 31(2):7, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho75].

Day:1976:CA

[Day76a] Samuel H. Day, Jr. Comrades in arms. Bulletin of the Atomic Scientists, 32(10):7, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SB77].

Day:1976:ELW

[Day76b] Samuel H. Day, Jr. Editorial: Lilienthal’s warning. Bulletin of the Atomic Scientists, 32(3):7, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1976:EPR

[Day76c] Samuel H. Day, Jr. Editorial: Plutonium recycling. Bulletin of the Atomic Scientists, 32(1):5, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1976:ERG

[Day76d] Samuel H. Day, Jr. Editorial: Remember Grand Forks. Bulletin of the Atomic Scientists, 32(4):5, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Dol76, You76b].

Day:1976:NNB

[Day76e] Samuel H. Day, Jr. Now or never for the B-1. Bulletin of the Atomic Scientists, 32(5):7, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1976:NB

[Day76f] Samuel H. Day, Jr. The nuclear button. Bulletin of the Atomic Scientists, 32(3):53, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1976:NG

[Day76g] Samuel H. Day, Jr. The numbers game. Bulletin of the Atomic Scientists, 32(8):6, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1976:SQA

[Day76h] Samuel H. Day, Jr. Some questions about Entebbe. Bulletin of the Atomic Scientists, 32(7):7, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [GRA^{+}76a, Lew76, Sho76, Ker76, Lev76, Dav76a, Ste76b, Bal76, Bar76b, Hor76, Ash76, Zar76, Pas76, Gin76, Wou76, Sla76, Tur76, Len76, Sol76, Med76, Hol76b, Whi77, Edm77].

Day:1977:ARC

[Day77a] Samuel H. Day, Jr. The arms race: a call to action. Bulletin of the Atomic Scientists, 33(9):6, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:BBR

[Day77b] Samuel H. Day, Jr. Books: Bloody reds, bloody yanks: Overkill, by John Cox; The Day Before Doomsday, by Sidney Lens. Bulletin of the Atomic Scientists, 33(7):63–64, September 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:CWL

[Day77c] Samuel H. Day, Jr. Converting the weapons labs. Bulletin of the Atomic Scientists, 33(1):27, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:CTP

[Day77d] Samuel H. Day, Jr. Correction: Toxicity of plutonium and some other actinides. Bulletin of the Atomic Scientists, 33(1):6, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds76].

Day:1977:ENC

[Day77e] Samuel H. Day, Jr. Editor’s note: [CERN]. Bulletin of the Atomic Scientists, 33(7):44, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:F

[Day77f] Samuel H. Day, Jr. A farewell. Bulletin of the Atomic Scientists, 33(10):4–5, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:FRD

[Day77g] Samuel H. Day, Jr. Foreword: Recombinant DNA research. Bulletin of the Atomic Scientists, 33(5):10, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:HKN

[Day77h] Samuel H. Day, Jr. Henry Kissinger: So near & yet so far. Bulletin of the Atomic Scientists, 33(2):4–5, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:LWF

[Day77i] Samuel H. Day, Jr. Lest we forget: the thoughts of three Nobel laureates on war and peace. Bulletin of the Atomic Scientists, 33(7):17, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:NDG

[Day77j] Samuel H. Day, Jr. Nuclear debate at the gut level. Bulletin of the Atomic Scientists, 33(3):56–58, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:NWL

[Day77k] Samuel H. Day, Jr. The nuclear weapons labs. Bulletin of the Atomic Scientists, 33(4):21–26, 28–32, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wel77].

Day:1977:OC

[Day77l] Samuel H. Day, Jr. Our contributors. Bulletin of the Atomic Scientists, 33(8):11, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:SC

[Day77m] Samuel H. Day, Jr. The science court. Bulletin of the Atomic Scientists, 33(4):43, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:SP

[Day77n] Samuel H. Day, Jr. Swords and plowshares. Bulletin of the Atomic Scientists, 33(9):4–5, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1977:TMF

[Day77o] Samuel H. Day, Jr. Toward a more fragrant world. Bulletin of the Atomic Scientists, 33(5):6–7, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1979:MBP

[Day79] Samuel H. Day, Jr. In the matter of The Progressive and the ‘secret’. Bulletin of the Atomic Scientists, 35(7):8–9, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1988:BRB

[Day88] Samuel H. Day, Jr. Book review: At Work in the Fields of the Bomb, by Robert Del Tredici. Bulletin of the Atomic Scientists, 44(8):46–47, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1992:PVI

[Day92] Samuel H. Day, Jr. Perspective: Vanunu: Israel’s embarrassment. Bulletin of the Atomic Scientists, 48(9):12–13, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BED93].

Day:1993:LCR

[Day93] Samuel H. Day, Jr. Letter: Costly revelation. Bulletin of the Atomic Scientists, 49(8):54, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coh93a].

Day:1995:PLW

[Day95] Samuel H. Day, Jr. Perspective: Letter from a Wisconsin jail. Bulletin of the Atomic Scientists, 51(5):16–17, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1996:BRN

[Day96] Samuel H. Day, Jr. Bulletin: Rotblat Nobel gives hope to Free-Vanunu campaign. Bulletin of the Atomic Scientists, 52(1):5– 6, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Day:1997:BRB

[Day97] Samuel H. Day, Jr. Book review: Breaking the code of silence: An Enemy of the State: The Life of Erwin Knoll, by Bill Lueders. Bulletin of the Atomic Scientists, 53(3):57–59, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-19368191.
html.

Daniel:1950:PAA

[DB50] Cuthbert Daniel and John L. Balderston. A proposal for an atomic armistice. Bulletin of the Atomic Scientists, 6(1):17, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dahl:1952:DCAa

[DB52a] Robert A. Dahl and Ralph S. Brown, Jr. Domestic control of atomic energy. [Part I]. Bulletin of the Atomic Scientists, 8(1): 17–22, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also parts II and III [DB52b, DB52c].

Dahl:1952:DCAb

[DB52b] Robert A. Dahl and Ralph S. Brown, Jr. Domestic control of atomic energy. Part II. Bulletin of the Atomic Scientists, 8(2): 56–59, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also parts I and III [DB52a, DB52c].

Dahl:1952:DCAc

[DB52c] Robert A. Dahl and Ralph S. Brown, Jr. Domestic control of atomic energy: Part III. Bulletin of the Atomic Scientists, 8(3): 88–91, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also parts I and II [DB52a, DB52b].

Durrani:1967:CRW

[DB67] S. A. Durrani and Gerald D. Berreman. Comments: On the role of women. Bulletin of the Atomic Scientists, 23(3):28–29, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber66b].

daCosta:1978:CUN

[dC78] João Frank da Costa. Commentary: From the United Nations: Instruments for development. Bulletin of the Atomic Scientists, 34(8):55–58, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:2014:ACC

[DC14] Frank W. Davis and Elizabeth A. Chornesky. Adapting to climate change in California. Bulletin of the Atomic Scientists, 70(5):62– 73, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Downey:1987:RK

[DCM87] Thomas J. Downey, Bob Carr, and Jim Moody. Report from Krasnoyarsk. Bulletin of the Atomic Scientists, 43(9):11–14, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davis:1979:CDV

[DD79] Donald E. Davis and Joe Dunphy. Commentary: Dioxin and the Vietnam veteran. Bulletin of the Atomic Scientists, 35(8): 58–59, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tho79b] and correction [Ano80h].

DeVolpi:1970:ES

[De 70a] Alexander De Volpi. Expectations from SALT. Bulletin of the Atomic Scientists, 26(4):6–8, 30–34, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeVolpi:1970:SRM

[De 70b] Alexander De Volpi. A special report: MIRV Gorgon Medusa of the Nuclear Age. Bulletin of the Atomic Scientists, 26(1):35–38, 46, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeVolpi:1974:EPD

[De 74] A. De Volpi. Energy policy decision-making: The need for balanced input. Bulletin of the Atomic Scientists, 30(10):29–33, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wil75g].

DeGroot:2005:BL

[De 05] Gerard J. De Groot. The bomb: a life. Harvard University Press, Cambridge, MA, USA, 2005. ISBN 0-674-01724-2. xiii + 397 + 16 pp. LCCN U264 .D43 2005.

Datt:1995:LGR

[DE95] Savita Datt and William Epstein. Letters: Getting real. Bulletin of the Atomic Scientists, 51(2):71–73, March/April 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eps94, Hol94f].

Diamond:2023:DCM

[DE23] Larry Diamond and James O. Ellis Jr. Deterring a Chinese military attack on Taiwan. Bulletin of the Atomic Scientists, 79 (2):65–71, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2178166.

Dean:1951:RAE

[Dea51] Gordon Dean. The role of atomic energy in the world economy. Bulletin of the Atomic Scientists, 7(6):185–188, June 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1952:AND

[Dea52] Gordon Dean. The atom in national defense. Bulletin of the Atomic Scientists, 8(8):289–290, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1953:CRB

[Dea53a] Gordon Dean. From chain reaction to breeding. Bulletin of the Atomic Scientists, 9(6):218, 226, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1953:NPD

[Dea53b] Gordon Dean. Nuclear power development: AEC policy statement. Bulletin of the Atomic Scientists, 9(8):314, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1954:HB

[Dea54a] Gordon Dean. The hydrogen bomb. Bulletin of the Atomic Scientists, 10(9):357, 362, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Strong negative review of the book The Hydrogen Bomb, by James Shepley and Clay Blair, Jr.

Dean:1954:TS

[Dea54b] Gordon Dean. Tasks for the statesmen. Bulletin of the Atomic Scientists, 10(1):11, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1983:FGA

[Dea83] Jonathan Dean. Federal Germany after the Euromissiles. Bulletin of the Atomic Scientists, 39(10):31–38, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1987:GAC

[Dea87] Jonathan Dean. Gorbachev arms control moves. Bulletin of the Atomic Scientists, 43(5):34–40, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1988:MED

[Dea88] Jonathan Dean. A menu of European defense plans. Bulletin of the Atomic Scientists, 44(7):23, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1989:CTG

[Dea89] Jonathan Dean. Conventional talks: a good first round. Bulletin of the Atomic Scientists, 45(8):24–31, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1991:NWO

[Dea91] Jonathan Dean. New world order: Coalitions for regional crises. Bulletin of the Atomic Scientists, 47(5):33–34, June 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1995:SUS

[Dea95] Jonathan Dean. A stronger U.N. strengthens America. Bulletin of the Atomic Scientists, 51(2):45–51, 53–54, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:1996:NNE

[Dea96] Jonathan Dean. No NATO expansion now. Bulletin of the Atomic Scientists, 52(3):18–19, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dean:2004:CC

[Dea04] Jonathan Dean. Counting casualties. Bulletin of the Atomic Scientists, 60(2):81–82, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/2/81.

DeBenedetti:1981:BRBb

[DeB81a] Charles DeBenedetti. Book review: A Turning Wheel: Three Decades of the Asian Revolution as Witnessed by a Correspondent for The New Yorker, by Robert Shaplen. Bulletin of the Atomic Scientists, 37(7):47–48, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeBenedetti:1981:BRBa

[DeB81b] Charles DeBenedetti. Book review: Fire in the Streets: America in the 1960s, by Milton Viorst. Bulletin of the Atomic Scientists, 37(5):46, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeCew:1948:NLR

[DeC48] Walter DeCew. New legislation to replace the McMahon Act. Bulletin of the Atomic Scientists, 4(9):277–279, 288, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1957:RFU

[Ded57] Stevan Dedijer. Research and freedom in undeveloped countries. Bulletin of the Atomic Scientists, 13(7):238–242, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1958:BDM

[Ded58] Steven Dedijer. The birth and death of a myth. Bulletin of the Atomic Scientists, 14(5):164–168, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1959:WRP

[Ded59] Stevan Dedijer. Windowshopping for a research policy. Bulletin of the Atomic Scientists, 15(9):367–371, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1960:BBA

[Ded60] Stevan Dedijer. Books: Living on the Atomic Volcano, by Ivan Supek. Bulletin of the Atomic Scientists, 16(7):301, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1962:BBN

[Ded62a] Stevan Dedijer. Books: Science and the New Nations, edited by Ruth Gruber. Bulletin of the Atomic Scientists, 18(2):35–36, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1962:RMP

[Ded62b] Stevan Dedijer. Research: The motor of progress. Bulletin of the Atomic Scientists, 18(6):4–7, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1965:STN

[Ded65] Stevan Dedijer. Soviets take a new look at science. Bulletin of the Atomic Scientists, 21(3):40–41, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1966:OU

[Ded66] Stevan Dedijer. The other U.S. Bulletin of the Atomic Scientists, 22(2):28–30, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1970:BDA

[Ded70] Stevan Dedijer. The brain drain: An age-old problem. Bulletin of the Atomic Scientists, 26(3):9–11, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dedijer:1978:WWP

[Ded78] S. Dedijer. Watching the watchmen: Parliaments and national intelligence services. Bulletin of the Atomic Scientists, 34(6):40– 43, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deer:1957:RCU

[Dee57] James W. Deer. Reader comment: The unavoidable international shelter race. Bulletin of the Atomic Scientists, 13(2):66–67, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deer:1959:WHC

[Dee59] James W. Deer. Whatever happened to civil defense? Bulletin of the Atomic Scientists, 15(6):266–267, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deFord:1951:LEI

[deF51] Miriam Allen deFord. Letter to the Editor: [impure science]. Bulletin of the Atomic Scientists, 7(9):287, September 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wyl51].

DeGrasse:1984:MSE

[DeG84] Robert W. DeGrasse, Jr. The military: shortchanging the economy. Bulletin of the Atomic Scientists, 40(5):37–40, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deitch:2005:KDS

[Dei05] Christian Deitch. Kyrgyzstan: Democracy stalled? Bulletin of the Atomic Scientists, 61(1):??, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/16.

DAC:1960:NPA

[Dem60] Democratic Advisory Council. A national peace agency. Bulletin of the Atomic Scientists, 16(2):74–76, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeMeo:1976:LRP

[DeM76] James DeMeo. Letter: Roots of the problem. Bulletin of the Atomic Scientists, 32(2):4, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pre75].

Demchak:2019:CDD

[Dem19] Chris C. Demchak. China: Determined to dominate cyberspace and AI. Bulletin of the Atomic Scientists, 75(3):99– 104, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1604857.

Dennis:1947:LEI

[Den47] Don Dennis. Letter to the Editor: Important omission in account of world government. Bulletin of the Atomic Scientists, 3(7):197, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Denney:1953:RI

[Den53] Reuel Denney. Reactors of the imagination. Bulletin of the Atomic Scientists, 9(6):206–210, 224, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeNike:1965:LES

[DeN65] L. Douglas DeNike. Letter to the Editor: Sonic booms. Bulletin of the Atomic Scientists, 21(9):25, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeNike:1974:BBT

[DeN74a] L. Douglas DeNike. Books: Nuclear Theft: Risks and Safeguards, by Mason Willrich and Theodore B. Taylor. Bulletin of the Atomic Scientists, 30(8):48–49, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeNike:1974:RM

[DeN74b] L. Douglas DeNike. Radioactive malevolence. Bulletin of the Atomic Scientists, 30(2):16–20, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Coh75b] and response [DeN75].

DeNike:1975:LR

[DeN75] L. Douglas DeNike. Letters: Response. Bulletin of the Atomic Scientists, 31(2):3, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DeN74b, Coh75b].

DeNike:1976:LRC

[DeN76a] L. Douglas DeNike. Letter: Reverse China syndrome. Bulletin of the Atomic Scientists, 32(1):3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75b].

DeNike:1976:LGT

[DeN76b] L. Douglas DeNike. Letter: The genetic threat. Bulletin of the Atomic Scientists, 32(8):4, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Denzer:1977:LCP

[Den77] Peter W. Denzer. Letter: Collective paranoia. Bulletin of the Atomic Scientists, 33(6):3, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeNike:1978:LAH

[DeN78] L. Douglas DeNike. Letter: Atomic haste. Bulletin of the Atomic Scientists, 34(3):2–3, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Densler:1982:JCD

[Den82] David Densler. “Just in case” — the danger of flexible response. Bulletin of the Atomic Scientists, 38(9):55–57, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dennett:2014:RCN

[Den14] Lydia Dennett. A realignment commission for national labs: How to downsize America’s bloated and unsecure nuclear weapons complex. Bulletin of the Atomic Scientists, 70(6):39–47, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dernberger:1966:ER

[Der66] Robert F. Dernberger. Economic realities. Bulletin of the Atomic Scientists, 22(6):6–10, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dernberger:1969:ERC

[Der69] Robert F. Dernberger. Economic realities and China’s political economics. Bulletin of the Atomic Scientists, 25(2):34–42, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Dertouzos:1985:LDU

[Der85] Michael L. Dertouzos. Letter: DARPA and U.S. technological future. Bulletin of the Atomic Scientists, 41(3):62, March 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [OSS84, CO85, Kal85b, Orn85].

Derber:1988:BRB

[Der88] Charles Derber. Book review: Beyond the Laboratory: Scientists as Political Activists in 1930s America, by Peter J. Kuznick. Bulletin of the Atomic Scientists, 44(5):52, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deutch:1960:CWA

[Deu60] Michael J. Deutch. Can we afford atomic power for underdeveloped countries? Bulletin of the Atomic Scientists, 16(1):23–27, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deutsch:1967:ACE

[Deu67] Karl W. Deutsch. Arms control and European unity: The next ten years. Bulletin of the Atomic Scientists, 23(5):21–24, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deutsch:1968:ASL

[Deu68] Steven Deutsch. Automation: A study of local union leaders. Bulletin of the Atomic Scientists, 24(5):52–53, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deudney:1983:BRBa

[Deu83a] Daniel Deudney. Book review: Confrontation in Space, by G. Harry Stine. Bulletin of the Atomic Scientists, 39(4):35–37, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deudney:1983:BRBb

[Deu83b] Daniel Deudney. Book reviews: High Frontier: A New National Strategy, by Daniel O. Graham; War in Space, by James Canan; Spacewar, by David Ritchie; The Shape of Wars to Come, by David Baker; Outer Space: A New Dimension of the Arms Race, edited by Bhupendra Jasani. Bulletin of the Atomic Scientists, 39(7):46–48, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deudney:1984:CAS

[Deu84a] Daniel Deudney. Commentary: Arms sale of the century. Bulletin of the Atomic Scientists, 40(1):57–58, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deudney:1984:CWG

[Deu84b] Daniel Deudney. Commentary: ‘What goes up must come down’. Bulletin of the Atomic Scientists, 40(3):10–11, March 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Deudney:1991:ESM

[Deu91] Daniel Deudney. Environment and security: Muddled thinking. Bulletin of the Atomic Scientists, 47(3):22–28, April 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bye91].

DeVries:1962:LEN

[DeV62] H. DeVries. Letter to the Editor: Non-nuclear club. Bulletin of the Atomic Scientists, 18(10):32, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Devereux:1969:BBP

[Dev69] Edward C. Devereux, Jr. Books: The Poverty of Liberalism, by Robert Paul Wolff; Violence and Social Change: A Review of Current Literature, by Henry Bienen. Bulletin of the Atomic Scientists, 25(10):36–38, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeVolpi:1988:LCI

[DeV88] Alex DeVolpi. Letter: On closer inspection. Bulletin of the Atomic Scientists, 44(2):52–53, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GK87].

DeVolpi:1992:LTT

[DeV92] Alex DeVolpi. Letter: Two-track verification for chemical treaty. Bulletin of the Atomic Scientists, 48(1):45–47, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi91].

DeVolpi:1995:FFP

[DeV95] Alex DeVolpi. Fast finish to plutonium peril. Bulletin of the Atomic Scientists, 51(5):20–21, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PWS^{+}96].

Dewey:1948:CAA

[Dew48] Thomas E. Dewey. The challenge of the Atomic Age. Bulletin of the Atomic Scientists, 4(10):296–297, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ano48-30, Ano48r, McM48a] and response [Tru48a].

Dews:1952:SPB

[Dew52] Edmund Dews. Science and politics in Britain: The Paymaster-General. Bulletin of the Atomic Scientists, 8(5):147–150, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dews:1953:TBR

[Dew53] Edmund Dews. Two British reports. Bulletin of the Atomic Scientists, 9(3):81–82, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pri53].

Dews:1955:BBO

[Dew55] Edmund Dews. Books: The Organization of Applied Research in Europe, the United States, and Canada, by the Organization for European Economic Cooperation. Bulletin of the Atomic Scientists, 11(2):64, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1960:SSI

[DeW60] Nicholas DeWitt. Soviet science: The institutional debate. Bulletin of the Atomic Scientists, 16(6):208–211, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWeerd:1962:ACL

[DeW62] Harvey A. DeWeerd. Arms control and the legacy of the past. Bulletin of the Atomic Scientists, 18(6):28–31, June 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1979:UGD

[DeW79] Hugh E. DeWitt. Has US government disclosed the secret of the H-bomb? Bulletin of the Atomic Scientists, 35(6):60–62, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1981:CMS

[DeW81] Hugh E. DeWitt. Commentary: On ’Military-scientists’. Bulletin of the Atomic Scientists, 37(8):60, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar81f].

DeWalt:1983:CEF

[DeW83] Billie R. DeWalt. The cattle are eating the forest. Bulletin of the Atomic Scientists, 39(1):18–23, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1984:LDA

[DeW84] Hugh E. DeWitt. Labs drive the arms race. Bulletin of the Atomic Scientists, 40(9):40–42, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Das85, May85].

DeWitt:1988:LXR

[DeW88] Hugh E. DeWitt. Letter: X-ray laser hype confirmed. Bulletin of the Atomic Scientists, 44(8):52, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu88].

deForest:1979:RTBf

[dF79a] Paul de Forest. The reading table: Arms Control and SALT II, by W. K. H. Panofsky; An Idea and Its Servants: UNESCO from Within, by Richard Hoggart. Bulletin of the Atomic Scientists, 35 (8):42, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBe

[dF79b] Paul de Forest. The reading table: Nuclear Policy and World Order: Why Denuclearization, by Richard A. Falk; Janus: A Summing Up, by Arthur Koestler. Bulletin of the Atomic Scientists, 35(7):50, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBc

[dF79c] Paul de Forest. The reading table: The Politics of Cancer, by Samuel S. Epstein; The Pesticide Conspiracy, by Robert van den Bosch. Bulletin of the Atomic Scientists, 35(5):41, May 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBh

[dF79d] Paul de Forest. The reading table: The Technological Conscience: Survival and Dignity in an Age of Expertise, by Manfred Stanley; The Limits of Technocratic Politics, by Jeffrey D. Straussman; The Irony of Vietnam: The System Worked, by Leslie H. Gelb and Richard K. Betts. Bulletin of the Atomic Scientists, 35 (10):42–43, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBd

[dF79e] Paul de Forest. The reading table: The Visible Scientists, by Rae Goodell; Scientists in Organizations: Productive Climates for Research and Development, by Donald C. Pelz and Frank M. Andrews. Bulletin of the Atomic Scientists, 35(6):57–58, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBa

[dF79f] Paul de Forest. The reading table: Toward a Metric of Science: the Advent of Science Indicators, edited by Yehuda Elkana, Joshua Lederberg, Robert K. Merton, Arnold Thackray, and Harriet Zuckerman; Reconciling Man with the Environment, by Eric Ashby; Science, Sin and Scholarship, edited by Irving Louis Horowitz; Scientists Confront Velikovsky, edited by Donald Goldsmith. Bulletin of the Atomic Scientists, 35(1):46–47, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBg

[dF79g] Paul de Forest. The reading table: War and the Liberal Conscience, by Michael Howard; The Endurance of Life, by MacFarlane Burnet; Science, Technology and Development: Essays in Honor of Professor A. Rahman, edited by K. D. Sharma and M. A. Quareshi; Paradoxes of Progress, by Gunther S. Stent. Bulletin of the Atomic Scientists, 35(9):55–56, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1979:RTBb

[dF79h] Paul de Forest. The reading table: William Henry Bragg, 1862– 1942: Man and Scientist, by G. M. Caroe; The U.S. Senate and Strategic Arms Policy, 1969–1977, by Alan Platt; No Other Gods: On Science and American Social Thought, by Charles E. Rosenberg; Atlantis: Fact or Fiction?, edited by Edwin S. Ramage. Bulletin of the Atomic Scientists, 35(4):38–39, April 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deForest:1980:RTB

[dF80] Paul de Forest. The reading table: Science Textbook Controversies and Politics of Equal Time, by Dorothy Nelkin; Controversy:

The Politics of Technical Decisions, edited by Dorothy Nelkin. Bulletin of the Atomic Scientists, 36(4):51, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Doering:1963:LEU

[DFF^{+}63] William Doering, Bernard T. Feld, Jr., Allan Forbes, James G. Patton, and Leo Szilard. Letter to the Editor: U.N. in Cuba. Bulletin of the Atomic Scientists, 19(4):33, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Davison:1963:LEP

[DFJ^{+}63] William C. Davison, Jerome D. Frank, Herbert Jehle, Steward Meacham, Clarence E. Pickett, Robert Fink, Kurt Kreith, and David R. Inglis. Letter to the Editor: The perils of deterrence. Bulletin of the Atomic Scientists, 19(3):28–30, March 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Delgado:1979:CSF

[DG79] Richard Delgado and Harold P. Green. Commentary: Science and the first amendment. Bulletin of the Atomic Scientists, 35(1):60– 62, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gre78].

Dumas:1986:ECE

[DG86] Lloyd J. Dumas and Suzanne Gordon. Economic conversion: an exchange. Bulletin of the Atomic Scientists, 42(6):45–48, June/ July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drell:2007:WNW

[DG07] Sidney D. (Sidney David) Drell and James E. Goodby. What are nuclear weapons for?: recommendations for restructuring U.S. strategic nuclear forces. Arms Control Association, Washington, DC, revised and updated edition, 2007. vi + 32 pp.

LCCN JZ5687 .D74 2007. URL http://armscontrol.org/pdf/
20071104_Drell_Goodby_07_new.pdf.

deHoffmann:1975:AOT

[dH75] Frederic de Hoffmann. ‘All in our time’: Pure science in the service of wartime technology. Bulletin of the Atomic Scientists, 31(1):41–44, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dorman:1986:CES

[DH86] William A. Dorman and Daniel Hirsch. Chernobyl: the emerging story: The U.S. media’s slant. Bulletin of the Atomic Scientists, 42(7):54–56, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Durrant:1990:DSB

[DH90] Damian Durrant and Joshua Handler. Deep-six B90 bomb. Bulletin of the Atomic Scientists, 46(5):20, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dorn:1998:LCD

[DH98] Greg Dorn and A. C. Hall. Letters: Cognitive dissonance. Bulletin of the Atomic Scientists, 54(1):67–68, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo97k].

Dhanapala:1990:NRD

[Dha90] Jayantha Dhanapala. NPT review: Disappointment in the Third World. Bulletin of the Atomic Scientists, 46(6):30–31, July/ August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dhara:2016:CCT

[Dha16] Sagar Dhara. Climate change, technological innovation: an Indian response. Bulletin of the Atomic Scientists, 72(1):7–9, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Diamond:1963:RS

[Dia63] Edwin Diamond. The rites of spring. Bulletin of the Atomic Scientists, 19(5):26–29, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Diamond:1983:BRB

[Dia83] Edwin Diamond. Book review: Thinking about the Next War, by Thomas Powers; With Enough Shovels: Reagan, Bush and Nuclear War, by Robert Scheer. Bulletin of the Atomic Scientists, 39(6):32, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Diamond:1984:BRB

[Dia84] Edwin Diamond. Book review: Open Skies, by Walter W. Rostow. Bulletin of the Atomic Scientists, 40(7):39–40, August/ September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dickerman:1977:RCC

[Dic77] R. L. Dickerman. Rebuttal: The case for continued laser uranium enrichment. Bulletin of the Atomic Scientists, 33(6):54–56, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cas77b, Cas77c, Gar77c].

Dickson:2001:SSC

[Dic01] Paul Dickson. Sputnik: the shock of the century. Walker and Company, 435 Hudson Street, New York, NY 10014, USA, 2001. ISBN 0-8027-1365-3 (ib.). 10 + 310 pp. LCCN ????

Dickson:2003:SSC

[Dic03] Paul Dickson. Sputnik: the shock of the century. Berkley Books, New York, NY, USA, 2003. ISBN 0-425-18843-4. 310 pp. LCCN TL796.5.S652.S664; TL796.5.S652.S664 2003.

Dickson:2011:SSC

[Dic11] Paul Dickson. Sputnik: the shock of the century. Walker and Company, New York, NY, USA, 2011. ISBN 0-8027-7951-4. 310 pp. LCCN TL796.5.S652 S664 2011. URL http://catdir.loc.

gov/catdir/enhancements/fy0612/2001026156-d.html.

Dietz:1991:LTD

[Die91] Leonard A. Dietz. Letter: Threat from depleted-uranium cannon shells. Bulletin of the Atomic Scientists, 47(2):47, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1973:PPP

[DIJ73] Carl Djerassi, Andrew Israel, and Wolfgang Jochle. Planned parenthood for pets? Bulletin of the Atomic Scientists, 29(1):10–19, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SB73] and response [Dje73].

Dillon:1957:WFB

[Dil57] A. C. Douglas Dillon. Why France and Britain called off their invasion of Egypt. Bulletin of the Atomic Scientists, 13(1):3, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dillon:1966:FIB

[Dil66] Wilton S. Dillon. The flow of ideas between Africa and America. Bulletin of the Atomic Scientists, 22(4):23–26, April 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dinell:1969:BBF

[Din69] Tom Dinell. Books: Academic Freedom and Tenure, edited by Louis Joughlin. Bulletin of the Atomic Scientists, 25(4):46–47, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kug69].

Dine:1978:MRC

[Din78] Thomas A. Dine. Military R&D: Congress’ next area of policy penetration. Bulletin of the Atomic Scientists, 34(2):28–37, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dismukes:1979:WSS

[Dis79] Key Dismukes. What should society expect from scientists? Bulletin of the Atomic Scientists, 35(9):19–21, November 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dishman:2004:BT

[Dis04] Chris D. Dishman. Bankrolling terror. Bulletin of the Atomic Scientists, 60(3):60–62, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/3/60.

Divine:1970:BBP

[Div70] Robert A. Divine. Books: The Politics of Peace-Keeping, by Alan James. Bulletin of the Atomic Scientists, 26(3):41, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Divine:1979:BRB

[Div79] Robert A. Divine. Book reviews: Congress and Arms Control, edited by Alan Platt and Lawrence D. Weiler; Negotiating Security: An Arms Control Reader, edited by William H. Kincade and Jeffrey D. Porro . Bulletin of the Atomic Scientists, 35(9):49– 50, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pie80, Kin80a].

Divine:1986:ERT

[Div86] Robert A. Divine. Early record on test moratoriums. Bulletin of the Atomic Scientists, 42(5):24–26, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Diwan:1979:BRB

[Diw79] Romesh Diwan. Book review: Technology and Underdevelopment, by Frances Stewart. Bulletin of the Atomic Scientists, 35(9):50, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dixon:1965:MHH

[Dix65] James P. Dixon. Man and his habitat: The polluted air: For air conservation. Bulletin of the Atomic Scientists, 21(6):6–12, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dixon:1966:LEE

[Dix66] John R. Dixon. Letter to the Editor: Engineering, affluence, and poverty. Bulletin of the Atomic Scientists, 22(4):32–33, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dizdarevic:1994:SS

[Diz94] Zlatko Dizdarevic. Sarajevo stories. Bulletin of the Atomic Scientists, 50(2):32–37, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deJouvenel:1957:PEP

[dJ57a] Bertrand de Jouvenel. From political economy to political ecology. Bulletin of the Atomic Scientists, 13(8):287–291, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deJouvenel:1957:CSE

[dJ57b] Bertrand de Jouvenel. On the character of the Soviet economy. Bulletin of the Atomic Scientists, 13(9):327–330, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Fow58].

deJouvenel:1963:PCR

[dJ63] Bertrand de Jouvenel. The politics consequences of the rise of science. Bulletin of the Atomic Scientists, 19(10):2–8, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [GF64].

deJouvenel:1964:LFT

[dJ64] Bertrand de Jouvenel. Letter from France: The Technocratic Age. Bulletin of the Atomic Scientists, 20(8):27–29, October 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Delucchi:2013:MWE

[DJ13] Mark A. Delucchi and Mark Z. Jacobson. Meeting the world’s energy needs entirely with wind, water, and solar power. Bulletin of the Atomic Scientists, 69(4):30–40, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/4/30.

Djerassi:1968:HPR

[Dje68] Carl Djerassi. A high priority? Research centers in developing nations. Bulletin of the Atomic Scientists, 24(1):22–27, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1972:CMC

[Dje72a] Carl Djerassi. Can man control his biological evolution? — a symposium on genetic engineering: V. probabilities and practicalities. Bulletin of the Atomic Scientists, 28(10):25–28, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1972:FCT

[Dje72b] Carl Djerassi. Fertility control through abortion: An assessment of the period 1950–1980. Bulletin of the Atomic Scientists, 28(1): 9–14, 41–45, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1973:CR

[Dje73] Carl Djerassi. Communications: Response. Bulletin of the Atomic Scientists, 29(7):54–55, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DIJ73, SB73].

Djerassi:1974:CAF

[Dje74] Carl Djerassi. The Chinese achievement in fertility control: Onethird of the women of child-bearing age may be practicing birth control. Bulletin of the Atomic Scientists, 30(6):17–25, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1975:ETP

[Dje75] Carl Djerassi. Editorial: Turning a powder keg into a fire hose. Bulletin of the Atomic Scientists, 31(6):9–10, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1976:MPI

[Dje76] Carl Djerassi. A modest proposal for increased north-south interaction among scientists. Bulletin of the Atomic Scientists, 32 (2):56–60, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Djerassi:1981:BCY

[Dje81] Carl Djerassi. Birth control in the year 2001. Bulletin of the Atomic Scientists, 37(3):24–28, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnay:1982:DP

[DJE^{+}82] Albert Donnay, Carl G. Jacobsen, R. L. Evans, Henry E. Eccles, and Colin Gray. Deterrence is the problem. Bulletin of the Atomic Scientists, 38(5):60–62, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra81b].

Djerassi:1986:PAU

[Dje86] Carl Djerassi. Population: Abortion in the United States: politics or policy? Bulletin of the Atomic Scientists, 42(4):38–41, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzenbach:1960:BBL

[dK60] Nicholas deB. Katzenbach. Books: Atoms and the Law, by E. Blyth Stason, Samuel D. Estep, and William J. Pierce. Bulletin of the Atomic Scientists, 16(8):341, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dollezhal:1980:NPE

[DK80] N. Dollezhal and Y. Koryakin. Nuclear power engineering in the Soviet Union. Bulletin of the Atomic Scientists, 36(1):33–37, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deKlerk:1983:CSS

[dKL83] Piet de Klerk and William Lanouette. Commentary: On ‘Safeguards Secrecy’. Bulletin of the Atomic Scientists, 39(9):62–63, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lan83b].

deLaguna:1959:WSW

[dL59] Wallace de Laguna. What is safe waste disposal? Bulletin of the Atomic Scientists, 15(1):35–43, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deLeon:1968:SMS

[dL68] Benjamin de Leon. Is science morally sterile? Bulletin of the Atomic Scientists, 24(5):54–55, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1978:TPB

[DL78] John Dowling and Milton Leitenberg. Television program: In the Event of Catastrophe. Bulletin of the Atomic Scientists, 34(4): 46–47, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:2021:LDE

[DL21] Jacqueline A. Dowling and Nathan S. Lewis. Long-duration energy storage for reliable renewable electricity: The realistic possibilities. Bulletin of the Atomic Scientists, 77(6):281–284, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1989191.

DeVolpi:1979:CMT

[DLF79] A. De Volpi, Amory B. Lovins, and Bernard T. Feld. Commentary: More on thorium cycles and nuclear proliferation. Bulletin of the Atomic Scientists, 35(7):57–59, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lov79].

DeGrasse:1981:HCR

[DM81a] Robert DeGrasse Jr. and Paul Murphy. The high costs of rearmament. Bulletin of the Atomic Scientists, 37(8):16–23, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1981:CSC

[DM81b] Hugh E. DeWitt and Gerald E. Marsh. Commentary: Secrecy and the comprehensive test ban. Bulletin of the Atomic Scientists, 37 (9):53–54, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dahlberg:1982:NGH

[DM82] Susan Dahlberg and Tom McNevin. The name of the game: Half-SAFE. Bulletin of the Atomic Scientists, 38(4):35–38, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ano82l] and erratum [Ano82l].

DeWitt:1984:CSR

[DM84] Hugh E. DeWitt and Gerald E. Marsh. Commentary: Stockpile reliability and nuclear testing. Bulletin of the Atomic Scientists, 40(4):40–41, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lar84].

DeWitt:1985:WDP

[DM85] Hugh E. DeWitt and Gerald E. Marsh. Weapons design policy impedes test ban. Bulletin of the Atomic Scientists, 41(10):10– 13, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1986:UTB

[DM86] Hugh E. DeWitt and Gerald E. Marsh. An update on the test ban. Bulletin of the Atomic Scientists, 42(4):10, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gar86, Mey86].

Dahlman:2005:SYP

[DM05] Ola Dahlman and Jenifer Mackby. Sealed for your protection. Bulletin of the Atomic Scientists, 61(4):76, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/4/76.

Diaz-Maurin:2018:CLT

[DM18] François Diaz-Maurin. Chronic long-term risk of low-level radiation exposure: Bridging the lay/expert divide. Bulletin of the Atomic Scientists, 74(5):335–339, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1507792.

Diaz-Maurin:2024:ERN

[DM24a] François Diaz-Maurin. The enduring risks and new challenges of nuclear materials: A special issue dedicated to Rodney C. Ewing’s scientific and policy contributions. Bulletin of the Atomic Scientists, 81(1):1–2, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2445937.

Diaz-Maurin:2024:HFR

[DM24b] François Diaz-Maurin. How Fukushima’s radioactive fallout in Tokyo was concealed from the public. Bulletin of the Atomic Scientists, 81(1):17–28, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2445939.

Diaz-Maurin:2024:NTC

[DM24c] François Diaz-Maurin. Nuclear testing in the 21st century — legacy, tensions, and risks. Bulletin of the Atomic Scientists, 80(2):73–74, 2024. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2024.2314438.

Dunning:1964:AOP

[DMD64] Gordon M. Dunning, Lindsay Mattison, and Richard Daly. AEC official protests. Bulletin of the Atomic Scientists, 20(7):29–30, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [MD64].

DeCarli:1998:LAC

[DMJ98] Paul S. DeCarli, Joseph J. Mangano, and Lefi Joslyn. Letters: Answering the challenge. Bulletin of the Atomic Scientists, 54 (4):66–67, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98].

Dahlman:2002:CB

[DMMH02] Ola Dahlman, Jenifer Mackby, Svein Mykkeltveit, and Hein Haak. Cheaters beware. Bulletin of the Atomic Scientists, 58(1): 28–35, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/1/28.

DuBridge:1967:CJS

[DN67] Lee A. DuBridge and H. L. Nieburg. Comments: The JPL story. Bulletin of the Atomic Scientists, 23(2):28, February 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nie66b].

deNevers:1971:AAP

[dN71] Noel de Nevers. Another approach to population control? Bulletin of the Atomic Scientists, 27(3):34, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deOyarzabal:1952:TGN

[dO52] Juan de Oyarzabal. The treatment of good neighbors. Bulletin of the Atomic Scientists, 8(7):251, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Dobzhansky:1949:SS

[Dob49] Theodosius Dobzhansky. The suppression of a science. Bulletin of the Atomic Scientists, 5(5):144–146, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dobzhansky:1952:SSR

[Dob52] Theodosius Dobzhansky. A symposium: The state of Russian science today: I: Lysenko’s “Michurinist” genetics. Bulletin of the Atomic Scientists, 8(2):40–44, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dobzhansky:1954:BBC

[Dob54] Th. Dobzhansky. Books: Origin of Cells From Living Matter and the Role of the Living Matter in the Organism, by O. B. Pepeshinskaia. Bulletin of the Atomic Scientists, 10(1):30, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dobzhansky:1964:EOS

[Dob64] Theodosius Dobzhansky. Evolution — organic and superorganic. Bulletin of the Atomic Scientists, 20(5):4–8, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Har64b].

Doctor:2007:CDM

[Doc07] Richard Doctor. Carbon dioxide on the move. Bulletin of the Atomic Scientists, 63(4):19–20, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/19.

Doherty:1957:BBI

[Doh57] Beka Doherty. Books: Gumption Island, by Felix Morley. Bulletin of the Atomic Scientists, 13(1):35, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Doherty:1958:BBB

[Doh58] Beka Doherty. Books: On the Beach, by Nevil Shute; The Genius and the Goddess, by Aldous Huxley. Bulletin of the Atomic Scientists, 14(2):93, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Doleman:1976:LAD

[Dol76] Edgar C. Doleman. Letter: ABM and defense. Bulletin of the Atomic Scientists, 32(7):6, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76d].

Dombey:1982:CIA

[Dom82] Norman Dombey. Commentary: International agreements on nuclear weapons. Bulletin of the Atomic Scientists, 38(3):36–39, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Domaratzki:1997:LJF

[Dom97] Z. Domaratzki. Letter: Just the facts. Bulletin of the Atomic Scientists, 53(2):61, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Arn96b].

Domenici:1998:DC

[Dom98] Pete V. Domenici. The Domenici challenge. Bulletin of the Atomic Scientists, 54(2):40–44, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [Pan98b, Lev98, Mar98b, TM98] and comments [Sai98, Mac98, DMJ98].

Donnay:1957:LEA

[Don57] J. D. H. Donnay. Letter to the Editor: Academic freedom in Italian universities. Bulletin of the Atomic Scientists, 13(4):152, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Boe57a, Boe57b].

Donnay:1964:LEP

[Don64] J. D. H. Donnay. Letter to the Editor: Pauling Prize. Bulletin of the Atomic Scientists, 20(3):25, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano63w].

Donnithorne:1966:CEC

[Don66] Audrey Donnithorne. Central economic control. Bulletin of the Atomic Scientists, 22(6):11–20, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnay:1982:CIC

[Don82] Albert Donnay. Commentary: On ‘inertial confinement’. Bulletin of the Atomic Scientists, 38(3):49, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Win81b].

Donnay:1985:LNF

[Don85] Albert Donnay. Letter: Nuclear-free zones. Bulletin of the Atomic Scientists, 41(9):54–55, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnelly:1986:BRBb

[Don86a] Warren H. Donnelly. Book review: Security or Armageddon, edited by Louis René Beres. Bulletin of the Atomic Scientists, 42(9):52, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnelly:1986:BRBa

[Don86b] Warren H. Donnelly. Book reviews: Safeguarding the Atom: A Critical Appraisal, by David Fischer and Paul Szasz; Non-Proliferation: The Why and the Wherefore, edited by Jozef Goldblat; The New Nuclear Nation, by Leonard S. Spector. Bulletin of the Atomic Scientists, 42(6):57–59, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnelly:1988:LMQ

[Don88] Warren H. Donnelly. Letter: Material question. Bulletin of the Atomic Scientists, 44(3):56–57, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack88n].

Donnelly:1989:BRB

[Don89] Warren H. Donnelly. Book review: The Nuclear Age: Atomic Energy, Proliferation and the Arms Race, second edition, by William Sweet. Bulletin of the Atomic Scientists, 45(7):41, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnelly:1990:BRE

[Don90] Warren H. Donnelly. Book review: Explosions for peace: Nuclear Dynamite: The Peaceful Nuclear Explosions Fiasco, by Trevor Findlay. Bulletin of the Atomic Scientists, 46(8):41–42, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Donnelly:1997:LK

[Don97] Warren H. Donnelly. Letter: Kudos. Bulletin of the Atomic Scientists, 53(4):59, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH97b].

Dooley:1959:SCF

[Doo59] D. J. Dooley. Science as cliché, fable, and faith. Bulletin of the Atomic Scientists, 15(9):372–375, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kos60, Hux59a, Aus60, Doo60].

Dooley:1960:LED

[Doo60] D. J. Dooley. Letter to the Editor: Dooley replies to Huxley. Bulletin of the Atomic Scientists, 16(1):33, January 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Doo59, Hux59a].

Doob:1964:LEG

[Doo64] Joseph L. Doob. Letter to the Editor: Government and university research. Bulletin of the Atomic Scientists, 20(4):33, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dorman:1984:JMC

[Dor84] William A. Dorman. Journalism and mass communications. Bulletin of the Atomic Scientists, 40(10):18s–19s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dorman:1985:MPG

[Dor85a] William A. Dorman. The media: playing the government’s game. Bulletin of the Atomic Scientists, 41(7):118–124, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pan86a].

Dorman:1985:SST

[Dor85b] William A. Dorman. Soviets seen through red-tinted glasses. Bulletin of the Atomic Scientists, 41(2):18–22, February 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dorn:1990:PUS

[Dor90] A. Walter Dorn. Perspective: U.N. should verify treaties. Bulletin of the Atomic Scientists, 46(6):12–13, July/August 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dornan:1996:LCP

[Dor96] Robert K. Dornan. Letter: Clinton policies destructive. Bulletin of the Atomic Scientists, 52(6):3, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa96e].

Dostrovsky:1972:WIN

[Dos72] Israel Dostrovsky. Water for Israel: New approaches to old problems. Bulletin of the Atomic Scientists, 28(8):11–18, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Doty:1965:FSD

[Dot65] Paul Doty. A freeze on strategic delivery systems. Bulletin of the Atomic Scientists, 21(2):2–6, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Douglas:1948:DC

[Dou48] William O. Douglas. Democracy and Communism. Bulletin of the Atomic Scientists, 4(5):135–137, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Douglas:1960:LEN

[Dou60] D. W. Douglas. Letter to the Editor: No cooperative atmosphere? Bulletin of the Atomic Scientists, 16(1):0, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dougherty:1968:NPP

[Dou68a] J. Chrys Dougherty. The narrow path to peace. Bulletin of the Atomic Scientists, 24(7):32–33, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Doumani:1968:RSP

[Dou68b] George Doumani. Reports: Science policy for Antarctica. Bulletin of the Atomic Scientists, 24(4):39–45, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dougherty:1969:CMC

[Dou69] J. Chrys Dougherty. Communiques: Military complex: The unpleasant symptom. Bulletin of the Atomic Scientists, 25(4):38, 42, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Douglas:2015:RST

[Dou15] Heather Douglas. Reshaping science: The trouble with the corporate model in Canadian government. Bulletin of the Atomic Scientists, 71(2):88–97, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Douglas:2024:CMP

[Dou24] Lawrence Douglas. Constitutional mistakes of the past can tyrannize the present — but we can fix them. Bulletin of the Atomic Scientists, 80(5):287–290, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2024.
2388463.

Dow:1967:CDA

[Dow67] M. C. G. McDonald Dow. Comments: Developing Africa. Bulletin of the Atomic Scientists, 23(8):42–43, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lar67].

Dowling:1977:NDF

[Dow77] John Dowling, Jr. The nuclear debate in film: Building of the Bomb vs. Lovejoy’s Nuclear War. Bulletin of the Atomic Scientists, 33(2):52–54, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gyo77].

Dowling:1978:CPA

[Dow78] John Dowling. Commentary: The public and the arms race. Bulletin of the Atomic Scientists, 34(10):47–49, December 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1979:FRBd

[Dow79a] John Dowling. Film review: Linus Pauling: Crusading Scientist, produced and directed by Robert Richter. Bulletin of the Atomic Scientists, 35(9):53, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1979:FRBa

[Dow79b] John Dowling. Film review: Nuclear Countdown, produced by Ramakantha Sarma. Bulletin of the Atomic Scientists, 35(2):48– 49, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1979:FRBe

[Dow79c] John Dowling. Film review: Paul Jacobs and the Nuclear Gang, produced by Jack Willis. Bulletin of the Atomic Scientists, 35 (10):44, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1979:FRBc

[Dow79d] John Dowling. Film review: The China Syndrome, produced by Michael Douglas and directed by James Bridges. Bulletin of the Atomic Scientists, 35(6):53–54, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1979:FRBb

[Dow79e] John Dowling. Film review: War Without Winners, produced by Harold Willens, and directed by Haskell Wexler. Bulletin of the Atomic Scientists, 35(4):33, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBe

[Dow80a] John Dowling. Film review: An American Ism: Joe McCarthy, produced and directed by Glen Silber. Bulletin of the Atomic Scientists, 36(5):57, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBf

[Dow80b] John Dowling. Film review: Boom! Bulletin of the Atomic Scientists, 36(6):57, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBd

[Dow80c] John Dowling. Film review: South Africa: The Nuclear File, produced by Peter Davis. Bulletin of the Atomic Scientists, 36 (5):57, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBg

[Dow80d] John Dowling. Film review: The Medical Implications of Nuclear Energy; Coal: Solution or Pollution?, produced by Ryall Wilson. Bulletin of the Atomic Scientists, 36(7):48, September 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBa

[Dow80e] John Dowling. Film review: We’ve Always Done It This Way; John, Mary, MIRV, and MaRV: The Arms Race and the Human Race; The End of the Rainbow, produced by Brian Kaufman; Reel Change: A Guide to Films on Appropriate Technology, by Diane Garey and Larry Hott. Bulletin of the Atomic Scientists, 36(2): 54, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBc

[Dow80f] John Dowling. Film reviews: A Play Half-Written — The Energy Adventure, produced by Bill Stokes; No Act of God, produced by Ian Ball and Sidney Goldsmith. Bulletin of the Atomic Scientists, 36(4):53–54, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBb

[Dow80g] John Dowling. Film reviews: An Act of Congress; Growth Dilemma, produced by Richard Whitacher and Christopher Aikenhead; Einstein, produced by Patrick Griffin. Bulletin of the Atomic Scientists, 36(3):55, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBh

[Dow80h] John Dowling. Film reviews: Manoevre, produced by Frederick Wiseman; Battleground Washington: The Politics of Pressure, produced by Phil Lewis; Hiroshima/Nagasaki, August 1945, produced by Erik Barnouw and Paul Ronder; Fable Safe, produced by Sumner Glimcher. Bulletin of the Atomic Scientists, 36(8): 55–56, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBj

[Dow80i] John Dowling. Film reviews: Sentenced to Success; Avital, produced by Mordechai Gal and Tal Larish; Clouds of Doubt; Reflections: Margaret Mead, produced by Tim White. Bulletin of the Atomic Scientists, 36(10):52–53, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1980:FRBi

[Dow80j] John Dowling. Film reviews: The Killing Ground, produced by Tom Priestly and Steve Singer; Here Today … Here Tomorrow. Radioactive Waste in America. Bulletin of the Atomic Scientists, 36(9):48, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:BRB

[Dow81a] John Dowling. Book review: The Genesis of New Weapons: Decision Making for Military R&D, edited by Franklin A. Long and Judith Reppy. Bulletin of the Atomic Scientists, 37(4):47–49, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBd

[Dow81b] John Dowling. Film review: Shadows of the Nuclear Age, produced by David Freudberg. Bulletin of the Atomic Scientists, 37 (7):49, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBf

[Dow81c] John Dowling. Film review: The Deadly Winds of War, produced by Michael Kendall. Bulletin of the Atomic Scientists, 37(8):42, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBg

[Dow81d] John Dowling. Film review: The Defense of the United States, a CBS Television Report in five parts: Ground Zero, The Nuclear Battlefield, Call to Arms, The War Machine, and The Russians. Bulletin of the Atomic Scientists, 37(9):46–48, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBe

[Dow81e] John Dowling. Film reviews: Focus on the Forties, directed by David Thaxton; The Life and Times of Rosie the Riveter, produced by Connie Field. Bulletin of the Atomic Scientists, 37(8): 41–42, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBb

[Dow81f] John Dowling. Film reviews: The Bull’s Eye War, produced by Dominic Flessati in 1976; The Deep Cold War, produced by John Penycate in 1977; The Real War in Space, produced by John Penycate in 1978; The Rise of the Red Navy, produced by Harry Hastings in 1978. Bulletin of the Atomic Scientists, 37 (5):47, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBa

[Dow81g] John Dowling. Film reviews: The MX Debate, produced by Shep Morgan; The Power Shift, produced by Ben Wattenberg; The SALT Syndrome; Second to None?, produced by Mike von Fremd; $1,000,000,000,000 for Defense, produced by Martin Koughan. Bulletin of the Atomic Scientists, 37(1):54–55, January 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBh

[Dow81h] John Dowling. Film reviews: The Red Army; The Last Epidemic: Medical Consequences of Nuclear Weaponry and Nuclear War. Bulletin of the Atomic Scientists, 37(10):36–37, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1981:FRBc

[Dow81i] John Dowling. Film reviews: Who’s in Charge Here, produced by Herb Goro; The Trials of Alger Hiss, produced by John Lowenthall. Bulletin of the Atomic Scientists, 37(6):47, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBb

[Dow82a] John Dowling. Film review: Acceptable Risk? The Nuclear Age in the United States. Bulletin of the Atomic Scientists, 38(2): 39, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBe

[Dow82b] John Dowling. Film review: Ecocide: A Strategy of War, produced by Tom Mayberry. Bulletin of the Atomic Scientists, 38 (5):54, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBc

[Dow82c] John Dowling. Film review: Eight Minutes to Midnight: A Portrait of Dr. Helen Caldicott, produced by Mary Benjamin. Bulletin of the Atomic Scientists, 38(3):35, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBa

[Dow82d] John Dowling. Film reviews: Ground Zero at Bangor, produced by J. Grayley Taylor; We Are the Guinea Pigs, produced by Ralph Klein and directed by Joan Harvey. Bulletin of the Atomic Scientists, 38(1):44, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBf

[Dow82e] John Dowling. Film reviews: Hiroshima: The People’s Legacy; To Die, To Live: The Survivors of Hiroshima; In Our Own Backyards: Uranium Mining in the United States, produced by Pamela Jones and Susanna Styron. Bulletin of the Atomic Scientists, 38 (8):49, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBd

[Dow82f] John Dowling. Film reviews: Nuclear War Graphics, produced under the guidance of Mike Casper; The Threat of Nuclear War, produced by Adam Suddaby; Return to Hiroshima, produced by Francis S. Lestingi. Bulletin of the Atomic Scientists, 38(4):44, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRBg

[Dow82g] John Dowling. Film reviews: Survivors, produced by Frances Politeo and Steven Okazaki; Thinking Twice About Nuclear War, produced by Kim Spencer and directed by Mark Centowski. Bulletin of the Atomic Scientists, 38(9):42, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1982:FRH

[Dow82h] John Dowling. Film reviews: How to run a film series on the arms race and nuclear war: Decision to Drop the Bomb; Fable Safe; Ground Zero; Hiroshima/Nagasaki August 1945; Last Epidemic; Nuclear Battlefield; Nuclear Countdown; $1,000,000,000,000 for Defense; Truman and the Atomic Bomb; The War Game; War Without Winners. Bulletin of the Atomic Scientists, 38(6):51–52, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Downey:1982:TSY

[Dow82i] Thomas J. Downey. Thirty-seven years of nuclear weapons: START … SALT … the freeze. Bulletin of the Atomic Scientists, 38(10):57–58, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1983:FRBd

[Dow83a] John Dowling. Film reviews: Hans Bethe: A Prophet of Energy, produced and distributed by Michael Blackwood; How Much is Enough: Decision Making in the Nuclear Age, produced by Andrew A. Stern. Bulletin of the Atomic Scientists, 39(5):44–45, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1983:FRBa

[Dow83b] John Dowling. Film reviews: If You Love This Planet: Dr. Helen Caldicott on Nuclear War, produced by Edward de Lorrain and directed by Terri Nash; Hiroshima/Nagasaki August 1945: The Case of the A-Bomb Footage; The Medical Consequences of Nuclear War. Bulletin of the Atomic Scientists, 39(1):39, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1983:FRBb

[Dow83c] John Dowling. Film reviews: Lovins on the Soft Path; Building the Brookhaven House; Kilowatts from Cowpies: The Methane Option, produced by Henry Mayer; Waterpower, produced by Neal Livingston. Bulletin of the Atomic Scientists, 39(3):44, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1983:FRBf

[Dow83d] John Dowling. Film reviews: No First Strike, produced and distributed by Union of Concerned Scientists; Nuclear War: A Guide to Armageddon, produced by Rick Green; War Without Winners II, produced by Haskell Wexler; Countdown for America. Bulletin of the Atomic Scientists, 39(9):47, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano83k], correcting No First Strike to No First Use.

Dowling:1983:FRBc

[Dow83e] John Dowling. Film reviews: No Place to Hide, produced and directed by Tom Johnson and Lance Bird; The Atomic Cafe, produced and directed by Kevin Rafferty, Pierce Rafferty, and Jayne Loader; Whose Budget Is It Anyway? Bulletin of the Atomic Scientists, 39(4):37, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1983:FRBe

[Dow83f] John Dowling. Film reviews: Picadon, produced by Renzo Kinoshita and distributed by Ron Santoni; One Million Hiroshimas, produced by Michael Anderson Films; Race to Oblivion; In the King of Prussia, produced by Emile de Antonio. Bulletin of the Atomic Scientists, 39(7):52, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Ano83b].

Dowling:1984:FRBb

[Dow84a] John Dowling. Film review: U.S. versus U.S.S.R: Who’s Ahead, produced by Gary Krane. Bulletin of the Atomic Scientists, 40 (3):50, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Ano84a].

Dowling:1984:FRBa

[Dow84b] John Dowling. Film reviews: Firebreaks: A War/Peace Game, produced and distributed by Ground Zero; Dark Circle; A Call for Survival, produced by Dan Grossman and Michael Roper; The Biology of Nuclear War. Bulletin of the Atomic Scientists, 40(1): 44, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1984:FRBc

[Dow84c] John Dowling. Film reviews: How Well We Meant; Facing Up to the Bomb; In Our Defense, produced by Bill Jersey and Michael Chandler. Bulletin of the Atomic Scientists, 40(4):37, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowling:1984:FRB

[Dow84d] John Dowling. Film reviews: Nuclear Strategy for Beginners, produced by Robert Zaiisk; What About the Russians?; Rings on the Water, produced and directed by Mercedes Gregory; Louder Than Our Words, directed by Lydia Dean Pilcher and Harriet Hirshorn. Bulletin of the Atomic Scientists, 40(6):42, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dow:1985:AFS

[Dow85] Michael McD. Dow. Africa: Food and security. Bulletin of the Atomic Scientists, 41(8):21–26, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dower:1986:BRB

[Dow86] John W. Dower. Book review: Japan’s Secret War, by Robert K. Wilcox. Bulletin of the Atomic Scientists, 42(7):61–62, August/ September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dower:1988:BRB

[Dow88] John W. Dower. Book review: Ideology and U.S. Foreign Policy, by Michael H. Hunt. Bulletin of the Atomic Scientists, 44(2): 48–50, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dower:1991:NWO

[Dow91] John W. Dower. New world order: Japan and the U.S. samurai spirit. Bulletin of the Atomic Scientists, 47(5):28–30, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dowdy:1992:LTC

[Dow92] Edward J. Dowdy. Letter: Technology centers for Russian scientists. Bulletin of the Atomic Scientists, 48(8):46, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap92].

Doyle:1953:LEC

[Doy53] William L. Doyle. Letter to the Editor: Current state of American visa policy. Bulletin of the Atomic Scientists, 9(1):21–22, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Doyle:2017:OOR

[Doy17] James E. Doyle. An occurrence at Oak Ridge: Morality in an age of nuclear peril. Bulletin of the Atomic Scientists, 73 (2):135–137, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1288456.

Doyle:2019:IND

[Doy19] James E. Doyle. The inhumanity of nuclear deterrence. Bulletin of the Atomic Scientists, 75(2):85–91, 2019. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1580893.

Doyle:2022:BNR

[Doy22] James E. Doyle. Building a nuclear off-ramp following the war in Ukraine. Bulletin of the Atomic Scientists, 78(4):218–223, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2087383.

Daniels:1953:TSC

[DP53] Farrington Daniels and Kenneth S. Pitzer. Two scientists comment. Bulletin of the Atomic Scientists, 9(5):164, June 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dorn:2012:CBH

[DP12] A. Walter Dorn and Robert Pauk. The closest brush: How a UN Secretary-General averted doomsday. Bulletin of the Atomic Scientists, 68(6):79–84, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/68/6/79.

Diringer:2020:CAG

[DP20] Elliot Diringer and Bob Perciasepe. The climate awakening of global capital. Bulletin of the Atomic Scientists, 76(5):233–237, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1806574.

Douglas:1975:LEF

[DPL75] John R. Douglas, Jr., David A. Patterson, and Amory B. Lovins. Letters: Energy for fertilizer. Bulletin of the Atomic Scientists, 31(5):2–3, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lov74b, Lov74c].

Davenport:2017:NDS

[DPT17] Kelsey Davenport, Jana Puglierin, and Petr Topychkanov. Nuclear disarmament summits: A proposal to break the international impasse. Bulletin of the Atomic Scientists, 73(4):264–270, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1338044.

Robin:1970:SLA

[dQR70] Gordon de Q. Robin. Science and logistics in Antarctica. Bulletin of the Atomic Scientists, 26(10):90–93, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

deRose:1949:AED

[dR49] François de Rose. The atomic energy debate at Paris: a French appraisal. Bulletin of the Atomic Scientists, 5(1):9–11, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Delcoigne:1971:NCP

[DR71] Georges C. Delcoigne and G. Rubinstein. Nonproliferation and control: Peaceful uses of atomic energy. Bulletin of the Atomic Scientists, 27(2):5–7, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dietz:1992:LNS

[DR92] Chris Dietz and Ernest A. Ryavec. Letter: “Nazi science” debate continued. Bulletin of the Atomic Scientists, 48(9):46, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GP92b].

Davis:2014:BEI

[DR14] Ethan B. Davis and Tom L. Richard. Biomass energy and the implications for climate and food: The US response. Bulletin of the Atomic Scientists, 70(1):16–20, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/70/1/16.full.pdf+html.

Drake:1969:BBR

[Dra69] Frank D. Drake. Books: Through Rugged Ways to the Stars, by Harlow Shapley. Bulletin of the Atomic Scientists, 25(10):

35, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drake:1973:CRE

[Dra73] Charles L. Drake. Conference reports: Earth and the oceans: The International Geodynamics Project. Bulletin of the Atomic Scientists, 29(8):32–33, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dragsdahl:2001:DD

[Dra01] Jergen Dragsdahl. The Danish dilemma. Bulletin of the Atomic Scientists, 57(5):45–50, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/45.

Dragonette:2006:ML

[Dra06] Charles N. Dragonette. Maritime legends. Bulletin of the Atomic Scientists, 62(5):18–19, September/October 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/5/18.

Dietz:1992:LA

[DRB92] Chris Dietz, Andrew Reynolds, and Gregory Benford. Letters: Agenda 2001. Bulletin of the Atomic Scientists, 48(7):44–45, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dreher:1946:LEW

[Dre46] Carl Dreher. Letters to the Editor: Wallace vs. Baruch. Bulletin of the Atomic Scientists, 2(9–10):30–31, November 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drechsler:1951:USD

[Dre51] Gerhard J. Drechsler. The U.S. State Department and world science. Bulletin of the Atomic Scientists, 7(4):121–122, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drell:1974:AAA

[Dre74] Sidney D. Drell. To act or not to act: ‘I have accepted the obligation to try to help the government function. Bulletin of the Atomic Scientists, 30(9):6, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bur74, Spe75].

Drell:1977:BSI

[Dre77] Sidney D. Drell. Beyond SALT II — a missile test quota. Bulletin of the Atomic Scientists, 33(5):34–42, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [RD77].

Drell:1982:FAC

[Dre82a] Sidney D. Drell. L + RV: a formula for arms control. Bulletin of the Atomic Scientists, 38(4):28–34, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dreyer:1982:BRB

[Dre82b] June Teufel Dreyer. Book review: Asian Security in the 1980s: Problems and Policies for a Time of Transition, edited by Richard H. Solomon. Bulletin of the Atomic Scientists, 38(3):34, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano82i].

Drell:1984:AR

[Dre84a] Sidney D. Drell. ABM revisited. Bulletin of the Atomic Scientists, 40(6):7–8, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dresbach:1984:CDF

[Dre84b] Robert L. Dresbach. Commentary: On ‘Downing of Flight 007’. Bulletin of the Atomic Scientists, 40(2):64–65, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tsi83a].

Drell:1987:GES

[Dre87] Sidney D. Drell. Guest editorial: Stop early SDI deployment. Bulletin of the Atomic Scientists, 43(3):3, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drell:1989:ACC

[Dre89a] Sidney Drell. Arms control criteria: Prevention first. Bulletin of the Atomic Scientists, 45(4):39, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drell:1989:F

[Dre89b] Sidney Drell. Not so fast. Bulletin of the Atomic Scientists, 45 (6):27–29, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drell:1991:SVT

[Dre91] Sidney D. Drell. START: Verification triumphs. Bulletin of the Atomic Scientists, 47(9):28–29, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dreyer:1999:TT

[Dre99] June Teufel Dreyer. Tangles up with Taiwan. Bulletin of the Atomic Scientists, 55(6):46–51, November/December 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/6/46.

Drell:2007:RPD

[Dre07] Sidney D. Drell. A reliable path to disarmament. Bulletin of the Atomic Scientists, 63(4):48–49, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/48.

Driscoll:1993:LHS

[Dri93] Jim Driscoll. Letter: Hawks in sheeps’ clothing. Bulletin of the Atomic Scientists, 49(4):47, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha93c, Isa93m].

Dennis:2009:SUS

[DRM^{+}09] Kate J. Dennis, Jason Rugolo, Lee T. Murray, Justin Parrella, David M. Romps, Christopher D. Holmes, Kurt Z. House, Benjamin G. Lee, and Mark T. Winkler. Should the United States resume reprocessing? A pro and con. Bulletin of the Atomic Scientists, 65(6):30–41, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/65/6/30.

deRose:1947:ARL

[dRMT^{+}47] François de Rose, General Andrew G. L. McNaughton, Sir George Thompson, Ignace Zlotowski, George H. Briggs, Edmundo de Holte Castello, and R. Gordon Arneson. Agency’s rights and limitations relating to inspections, surveys and explorations. Bulletin of the Atomic Scientists, 3(9):262–265, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drollette:2016:TSS

[Dro16a] Dan Drollette, Jr. Taking stock: Steven Chu, former Secretary of the Energy Department, on fracking, renewables, nuclear weapons, and his work, post-Nobel Prize. Bulletin of the Atomic Scientists, 72(6):351–358, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drollette:2016:VIP

[Dro16b] Dan Drollette, Jr. View from the inside: Prince Turki al-Faisal on Saudi Arabia, nuclear energy and weapons, and Middle East politics. Bulletin of the Atomic Scientists, 72(1):16–24, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drollette:2017:PWA

[Dro17] Dan Drollette, Jr. Pulitzer-winning author Tracy Kidder: Looking for the soul of the machine makers. Bulletin of the Atomic Scientists, 73(2):74–79, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1288433.

Drollette:2019:WIA

[Dro19] Dan Drollette Jr. What if the Arctic melts, and we lose the great white shield? interview with environmental policy expert Durwood Zaelke. Bulletin of the Atomic Scientists, 75 (5):239–246, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1654269.

Drollette:2020:PDB

[Dro20a] Dan Drollette, Jr. Peter Davis of the British Antarctic Survey on changes in the Thwaites Glacier. Bulletin of the Atomic Scientists, 76(3):121–128, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1751951.

Drollette:2020:SWW

[Dro20b] Dan Drollette Jr. Shorter, warmer winters, less snow. what next? Q&A with biologist Pamela Templer. Bulletin of the Atomic Scientists, 76(3):133–139, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1751966.

Drollette:2021:FTM

[Dro21a] Dan Drollette Jr. The five things that must happen for renewables to fit into the grid: Interview with Greg Nemet. Bulletin of the Atomic Scientists, 77(6):307–311, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1989206.

Drollette:2021:ICA

[Dro21b] Dan Drollette, Jr. Interview: CalPERS’ Anne Simpson on the climate change power of investment managers. Bulletin of the Atomic Scientists, 77(2):76–81, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2021.
1885846.

Drollette:2021:OWP

[Dro21c] Dan Drollette Jr. Offshore wind: Poised for the big time. An interview with Anthony Kirincich. Bulletin of the Atomic Scientists, 77(6):301–306, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1989202.

Drollette:2021:TSC

[Dro21d] Dan Drollette, Jr. Tom Steyer on clean energy: It’s where the big money is going. Bulletin of the Atomic Scientists, 77 (2):70–75, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1886445.

Drollette:2022:HTS

[Dro22a] Dan Drollette Jr. The high-tech surveillance state is not restricted to China: Interview with Maya Wang of Human Rights Watch. Bulletin of the Atomic Scientists, 78(5):239–242, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2109319.

Drollette:2022:ICW

[Dro22b] Dan Drollette Jr. Introduction: Can we grow and burn our way out of climate change? Bulletin of the Atomic Scientists, 78(3):123–124, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2066808.

Drollette:2022:IBN

[Dro22c] Dan Drollette Jr. Introduction: The brave new world of the high-tech surveillance state. Bulletin of the Atomic Scientists, 78 (5):237–238, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2109318.

Drollette:2022:IUU

[Dro22d] Dan Drollette Jr. Introduction: The unintended-and undermanagedconsequences of blockchain and cryptocurrency. Bulletin of the Atomic Scientists, 78(4):185–186, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2022.
2087366.

Drollette:2022:IWS

[Dro22e] Dan Drollette Jr. Introduction: Why some renewable technologies will perish in — and others survive — the “Valley of Death”. Bulletin of the Atomic Scientists, 79(1):1–2, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2157109.

Drollette:2022:SDC

[Dro22f] Dan Drollette Jr. Smart devices, cell phone cameras, social shaming and the loss of the right to a private self: Interview with Michel Paradis about the modern panopticon. Bulletin of the Atomic Scientists, 78(5):243–248, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2109322.

Drollette:2022:WDO

[Dro22g] Dan Drollette Jr. What do ordinary Russians think? Interview with a Russian independent reporter. Bulletin of the Atomic Scientists, 78(6):315–317, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2132737.

Drollette:2022:WIC

[Dro22h] Dan Drollette Jr. “When it comes to Russia, it’s like living in a volcano”: An interview with Farida Rustamova, an independent reporter working in Putin’s Russia. Bulletin of the Atomic Scientists, 78(5):268–272, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2109339.

Drollette:2022:WBC

[Dro22i] Dan Drollette Jr. Wood-burning: Carbon hero or carbon villain. Q&A with forest modeling scientist Michael Ter-Mikaelian. Bulletin of the Atomic Scientists, 78(3):158–161, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2065061.

Drollette:2023:CAS

[Dro23a] Dan Drollette Jr. Charging ahead: Steven Chu, Nobel Prizewinner and former Energy Secretary, on today’s battery research — and more. Bulletin of the Atomic Scientists, 79(6):366–371, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2266938.

Drollette:2023:DSO

[Dro23b] Dan Drollette, Jr. “He did not speak the ordinary language”: Memories of Oppie from a Manhattan Project physicist. Bulletin of the Atomic Scientists, 79(4):221–231, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2223075.

Drollette:2023:IES

[Dro23c] Dan Drollette Jr. Interview with Eric Schlosser: Why we can’t trust the government’s figures about nuclear close calls. Bulletin of the Atomic Scientists, 79(3):148–154, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2202535.

Drollette:2023:ISR

[Dro23d] Dan Drollette Jr. Interview with Sneha Revanur, “the Greta Thunberg of AI”. Bulletin of the Atomic Scientists, 80(1):3–8, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2293565.

Drollette:2023:ISS

[Dro23e] Dan Drollette Jr. Interview with Susan Solomon: The healing of the ozone hole, and what else we can learn from atmospheric near-misses. Bulletin of the Atomic Scientists, 79 (3):142–147, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2199567.

Drollette:2023:ICC

[Dro23f] Dan Drollette Jr. Introduction: Climate change-where are we now? Bulletin of the Atomic Scientists, 79(6):347, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2268398.

Drollette:2023:IHN

[Dro23g] Dan Drollette, Jr. Introduction: how to negotiate the China– Taiwan impasse. Bulletin of the Atomic Scientists, 79(2):55–56, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2179787.

Drollette:2023:INM

[Dro23h] Dan Drollette Jr. Introduction: Near-misses, close calls, and early warnings. Bulletin of the Atomic Scientists, 79(3):135–136, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2202526.

Drollette:2023:IWY

[Dro23i] Dan Drollette Jr. Introduction: What you can do to turn back the hands of the Clock. Bulletin of the Atomic Scientists, 80(1):1–2, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2295158.

Drollette:2023:LWB

[Dro23j] Dan Drollette Jr. “Like writing the biography of a ghost” — interview with Jeff goodell, author of The Heat Will Kill You First. Bulletin of the Atomic Scientists, 79(6):348–352, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2266927.

Drollette:2023:OVM

[Dro23k] Dan Drollette, Jr. Oppenheimer — “A very mysterious and delphic character.” Interview with Kai Bird, author of American Prometheus. Bulletin of the Atomic Scientists, 79(4):210–215, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2223071.

Drollette:2023:OMB

[Dro23l] Dan Drollette, Jr. Oppenheimer: The man behind the movie. Bulletin of the Atomic Scientists, 79(4):201–202, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2224160.

Drollette:2023:WCJ

[Dro23m] Dan Drollette Jr. Where climate journalism is now: Interview with Emily Atkin, the fire behind the Heated Climate newsletter. Bulletin of the Atomic Scientists, 79(6):360–365, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2266937.

Drollette:2024:AIW

[Dro24a] Dan Drollette Jr. After ITER: What China and others are doing in fusion. Interview with MIT’s Dennis Whyte. Bulletin of the Atomic Scientists, 80(6):371–376, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2418703.

Drollette:2024:FTA

[Dro24b] Dan Drollette Jr. Ferreting out the truth about fusion: Interview with Bob Rosner. Bulletin of the Atomic Scientists, 80 (6):345–351, 2024. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2024.2416832.

Drollette:2024:FMR

[Dro24c] Dan Drollette Jr. Figuring out the most realistic projections for sea-level rise: Interview with glaciologist Rob De-Conto. Bulletin of the Atomic Scientists, 80(4):218–224, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2364500.

Drollette:2024:FTB

[Dro24d] Dan Drollette Jr. “Fusion is not a typical bet.” Interview with Silicon Valley venture capitalist Mark Coopersmith. Bulletin of the Atomic Scientists, 80(6):358–364, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2416838.

Drollette:2024:HHS

[Dro24e] Dan Drollette Jr. “H is for hope” sounded a lot better than “D is for despair”: Interview with Elizabeth Kolbert about climate change. Bulletin of the Atomic Scientists, 80(4):230–234, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2364510.

Drollette:2024:IFN

[Dro24f] Dan Drollette Jr. Introduction: Fusion, the next big thing — again? Bulletin of the Atomic Scientists, 80(6):343–344, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2419714.

Drollette:2024:PIS

[Dro24g] Dan Drollette Jr. Praying for the ice (and snow, and water) as the climate changes. Bulletin of the Atomic Scientists, 80 (4):211–212, 2024. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2024.2365541.

Drollette:2025:CNT

[Dro25a] Dan Drollette, Jr. Carlos Nobre on tipping points in the Amazon rainforest. Bulletin of the Atomic Scientists, 81(2):97–100, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2025.2464432.

Drollette:2025:AHT

[Dro25b] Dan Drollette, Jr. Is the AMOC headed for a tipping point? Interview with Henk Dijkstra. Bulletin of the Atomic Scientists, 81(2):83–87, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2025.2465047.

Drollette:2025:TKI

[Dro25c] Dan Drollette, Jr. (Trying to) keep it cool: Vladimir Romanovsky on permafrost fieldwork. Bulletin of the Atomic Scientists, 81 (2):92–96, 2025. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2025.2464431.

Drucker:1971:IEE

[Dru71] Daniel C. Drucker. III. The engineer in the establishment. Bulletin of the Atomic Scientists, 27(10):31–34, December 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lev72b].

Drucker:1972:CEE

[Dru72] Daniel C. Drucker. Communications: On engineer in establishment. Bulletin of the Atomic Scientists, 28(5):4, May 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev72b].

Rose:1947:ODF

[dRVT^{+}47] François de Rose, John E. Vance, Sir George Thompson, General Andrew G. L. McNaughton, and Hsioh ren Wei. Operational and developmental functions of the international agency. Bulletin of the Atomic Scientists, 3(9):253–255, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniel:1947:FDR

[DS47a] Cuthbert Daniel and Arthur M. Squires. Freedom demands responsibility. Bulletin of the Atomic Scientists, 3(12):300–304, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ash48, Rit48, Ler48, Ano48v] and further responses [Ure48c, vL48b, Ure48c, Mas48b, DS49].

Daniel:1947:ICS

[DS47b] Cuthbert Daniel and Arthur M. Squires. The international control of safe atomic energy: Two engineers propose a compromise whereby the benefits of nuclear disarmament may outweigh the risks. Bulletin of the Atomic Scientists, 3(4–5):111–116, 135, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Daniel:1949:SRW

[DS49] Cuthbert Daniel and Arthur M. Squires. Scientists’ responsibilities on the way to peace, and after. Bulletin of the Atomic Scientists, 5(1):27–28, 32, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See original article [DS47a].

Doyle:2012:STI

[DS12] James Doyle and Charles Streeper. Steps toward increased nuclear transparency. Bulletin of the Atomic Scientists, 68(2): 55–62, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/2/55.

Douglas:2015:CCW

[DS15] Karen M. Douglas and Robbie M. Sutton. Climate change: Why the conspiracy theories are dangerous. Bulletin of the Atomic Scientists, 71(2):98–106, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dicks:1987:POF

[DSL^{+}87] Norm Dicks, Al Swift, Mike Lowry, Peter DeFazio, and Les Au-Coin. Politicians object to “FEMA’s blackmail plan”. Bulletin of the Atomic Scientists, 43(4):45, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pool:1960:VR

[dSP60] Ithiel da Sola Pool. A voice of realism? Bulletin of the Atomic Scientists, 16(1):6, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Three-paragraph excerpt from a review in the 4 December 1959 issue of Science.

Price:1965:SS

[dSP65] Derek J. de Solla Price. The science of science. Bulletin of the Atomic Scientists, 21(8):2–8, October 1965. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pas66].

Dreschhoff:1974:IHL

[DSZ74] Gisela Dreschhoff, D. F. Saunders, and E. J. Zeller. International high level nuclear waste management. Bulletin of the Atomic Scientists, 30(1):28–33, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drell:1991:ADT

[DT91] Sidney Drell and Charles H. Townes. Authors disagrees over test ban. Bulletin of the Atomic Scientists, 47(3):38–39, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deLanversin:2024:NCB

[dTdLF24] Julien de Troullioud de Lanversin and Christopher Fichtlscherer. New confidence-building measures can reduce tensions around subcritical tests. Bulletin of the Atomic Scientists, 80(2):94– 101, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00963402.2024.2314435.

delTredici:1993:FPT

[dTM93] Robert del Tredici and Mike Moore. First, puzzlement; then action. Bulletin of the Atomic Scientists, 49(2):24–29, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Duarte:1978:LEC

[Dua78] C. Duarte. Letter: Eyes closed. Bulletin of the Atomic Scientists, 34(4):3–4, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Swe77].

DuBridge:1946:RLL

[DuB46a] Lee A. DuBridge. The role of large laboratories in nuclear research. Bulletin of the Atomic Scientists, 2(9–10):12–13, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1946:SNP

[DuB46b] Lee A. DuBridge. Science and national policy. Bulletin of the Atomic Scientists, 1(11):12–14, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1946:WAB

[DuB46c] Lee A. DuBridge. What about the Bikini tests? Bulletin of the Atomic Scientists, 1(11):7, 16, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1948:ADP

[DuB48] Lee A. DuBridge. Assumption of duties is a personal matter. Bulletin of the Atomic Scientists, 4(3):74, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bri48a].

DuBridge:1949:WAC

[DuB49] Lee A. DuBridge. Why another cyclotron? Bulletin of the Atomic Scientists, 5(6–7):201–203, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubridge:1950:ENS

[Dub50a] Lee A. Dubridge. Editorial: National Science Foundation Act. Bulletin of the Atomic Scientists, 6(6):162, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1950:SF

[DuB50b] Lee A. DuBridge. The Science Foundation. Bulletin of the Atomic Scientists, 6(2):35–36, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubarle:1955:ORB

[Dub55a] Daniel Dubarle, O.P. Observations in the relations between science and state. Bulletin of the Atomic Scientists, 11(4):141–144, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

DuBridge:1955:WSR

[DuB55b] L. A. DuBridge. What is a security risk? Bulletin of the Atomic Scientists, 11(4):163, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Dubarle:1956:FRB

[Dub56a] Dominique Dubarle. The future of the relationship between science and religion. Bulletin of the Atomic Scientists, 12(10): 365–369, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cal57] and response [Dub57a].

Dubarle:1956:PIS

[Dub56b] Daniel Dubarle, O.P. Part III: Science and social responsibility: The scientist and his responsibilities. Bulletin of the Atomic Scientists, 12(7):253–257, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubarle:1957:LEG

[Dub57a] D. Dubarle. Letter to the Editor: Galileo rescinded? Bulletin of the Atomic Scientists, 13(5):184–185, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cal57, Dub56a].

DuBridge:1957:SEA

[DuB57b] Lee A. DuBridge. Science — the endless adventure. Bulletin of the Atomic Scientists, 13(3):74–79, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubarle:1959:TWC

[Dub59] Dominic Dubarle. Toward a world community of scientists. Bulletin of the Atomic Scientists, 15(5):178–180, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubinin:1960:PCW

[Dub60] M. M. Dubinin. Potentialities of chemical warfare. Bulletin of the Atomic Scientists, 16(6):250–251, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubinin:1962:ESI

[Dub62a] M. M. Dubinin. Exchanging scientific information. Bulletin of the Atomic Scientists, 18(8):13–15, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1962:FNC

[DuB62b] Lee A. DuBridge. [The first nuclear chain reaction]. Bulletin of the Atomic Scientists, 18(10):8, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1963:OMI

[DuB63] Lee A. DuBridge. [Our most important objectives in the space program]. Bulletin of the Atomic Scientists, 19(5):29, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1966:GRS

[DuB66] Lee A. DuBridge. The government role in science education. Bulletin of the Atomic Scientists, 22(5):16–20, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1969:CSC

[DuB69a] Lee A. DuBridge. Chicago: The social control of science and its applications. Bulletin of the Atomic Scientists, 25(5):26–28, 35, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DuBridge:1969:FUR

[DuB69b] Lee A. DuBridge. The future of university research. Bulletin of the Atomic Scientists, 25(1):39, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubos:1970:HL

[Dub70] René Dubos. The human landscape. Bulletin of the Atomic Scientists, 26(3):31–37, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dubey:1985:DMS

[Dub85] Muchkund Dubey. Deterrence masks superpower hegemony. Bulletin of the Atomic Scientists, 41(2):28–30, February 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dudley:1975:CMH

[Dud75a] H. C. Dudley. A commentary: Michelson’s hunch was right. Bulletin of the Atomic Scientists, 31(1):47–48, January 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil74b].

Dudley:1975:UCT

[Dud75b] H. C. Dudley. The ultimate catastrophe. Bulletin of the Atomic Scientists, 31(9):21–24, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and reply [vHWD76, DeN76a, Gla76b, Bet76].

Dudley:1976:YAT

[Dud76] H. C. Dudley. Are your assumptions tenable? Bulletin of the Atomic Scientists, 32(6):38, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet76].

Dudley:1977:LCS

[Dud77a] H. C. Dudley. Letter: Censorship by selection. Bulletin of the Atomic Scientists, 33(3):3–4, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ham76].

Dudley:1977:LNA

[Dud77b] Richard M. Dudley. Letter: Nuclear absolutes. Bulletin of the Atomic Scientists, 33(6):5, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel77n].

Duelfer:2021:SLT

[Due21] Charles A. Duelfer. Some long-term effects of UNSCOM: People are important, or, therein lies much of the problem. Bulletin of the Atomic Scientists, 77(4):188–191, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1941585.

Duffield:1970:ARN

[Duf70] Robert B. Duffield. An addendum [to Reflections on National Laboratories]. Bulletin of the Atomic Scientists, 26(2):15–16, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor70b].

Duffy:1986:ARS

[Duf86] Gloria Duffy. Administration redefines Soviet “violations”. Bulletin of the Atomic Scientists, 42(2):13–17, February 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Duffy:1987:SFT

[Duf87] Gloria Duffy. Study finds treaty compliance. Bulletin of the Atomic Scientists, 43(8):30–32, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dugger:1991:BRC

[Dug91] Ronnie Dugger. Book review: Campaign of error: Pledging Allegiance: The Last Campaign of the Cold War, by Sidney Blumenthal. Bulletin of the Atomic Scientists, 47(6):40–41, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dulles:1957:DP

[Dul57] John Foster Dulles. Disarmament and peace. Bulletin of the Atomic Scientists, 13(7):262–263, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dumas:1975:LMM

[Dum75] Lloyd J. Dumas. Letter: The military mind. Bulletin of the Atomic Scientists, 31(10):2, 4, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75f].

Dumas:1976:NSN

[Dum76] Lloyd J. Dumas. National security in the nuclear age. Bulletin of the Atomic Scientists, 32(5):24–35, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from The Papers of the Peace Society [Ano76c]. See [Hay76].

Dumas:1980:HFW

[Dum80] Lloyd J. Dumas. Human fallibility and weapons. Bulletin of the Atomic Scientists, 36(9):15–20, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dumas:1986:MBE

[Dum86] Lloyd J. Dumas. The military burden on the economy. Bulletin of the Atomic Scientists, 42(8):22–26, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SD87, Vog87].

Dumas:1993:SOC

[Dum93] Lloyd J. Dumas. Sanctions: Organizing the chaos. Bulletin of the Atomic Scientists, 49(9):46–49, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dunn:1948:SPR

[Dun48] L. C. Dunn. Science and politics in Russia. Bulletin of the Atomic Scientists, 4(12):368, 383, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dunn:1949:CAG

[Dun49] L. C. Dunn. Commentary by American geneticists: Motives for the purge. Bulletin of the Atomic Scientists, 5(5):142–143, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Duncan:1966:DE

[Dun66] James S. Duncan. The developing economy. Bulletin of the Atomic Scientists, 22(6):84–87, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dunham:1974:LAR

[Dun74] Joan Bixby Dunham. Letters: An angry reader. Bulletin of the Atomic Scientists, 30(6):4–5, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Edm74].

Dunbar:1975:LQ

[Dun75] Carl O. Dunbar. Letters: A question. Bulletin of the Atomic Scientists, 31(5):4, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kol75, Wil75a].

Dunlap:1983:CPC

[Dun83] Riley E. Dunlap. Commentary: Paradigm conflict. Bulletin of the Atomic Scientists, 39(9):59–61, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dunathan:1984:IIP

[Dun84] Harmon C. Dunathan. Inter-institution programs. Bulletin of the Atomic Scientists, 40(10):25s–26s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Duncan:1985:PHM

[Dun85] Jeffrey S. Duncan. Perspective: How many Soviet tests make a flurry? Bulletin of the Atomic Scientists, 41(9):8–9, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kal85c].

Dunn:1990:NRI

[Dun90] Lewis A. Dunn. NPT review: It ain’t broke — don’t fix it. Bulletin of the Atomic Scientists, 46(6):19–20, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Durand:1969:LES

[Dur69] Dorothy Durand. Letter to the Editor: Science for peace. Bulletin of the Atomic Scientists, 25(9):47, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dux:1952:NN

[Dux52] Dieter Dux. News and notes. Bulletin of the Atomic Scientists, 8(7):259–261, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

deVaucouleurs:1961:MP

[dV61] Gerard de Vaucouleurs. The Moon and planets. Bulletin of the Atomic Scientists, 17(5–6):214–217, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deVolpi:1978:CSP

[dV78] Alex de Volpi. Commentary: Soft on proliferation? Bulletin of the Atomic Scientists, 34(7):62–63, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano78c].

Drell:1991:LWP

[DvH91] Sidney D. Drell and Frank von Hippel. Letters: What panelist say about test ban, when they said it. Bulletin of the Atomic Scientists, 47(5):44–45, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH91].

Dvorin:1965:CCE

[Dvo65] Eugene P. Dvorin. The Chile–California experiment. Bulletin of the Atomic Scientists, 21(9):35–38, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dvorkin:2007:RRP

[Dvo07] Vladimir Dvorkin. Reading Russia’s posture. Bulletin of the Atomic Scientists, 63(4):16–18, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/16.

Dharma-wardana:1977:SHT

[Dw77] M. W. C. Dharma-wardana. Self-help for third world scientists. Bulletin of the Atomic Scientists, 33(2):22–23, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeVolpi:1985:CEP

[DW85] Alex DeVolpi and Dee Wernette. Congress enacts a peace institute. Bulletin of the Atomic Scientists, 41(2):42, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

deWaard:1990:LSW

[dW90a] Anita de Waard. Letter: And Stalin wrote back to Kapitsa. Bulletin of the Atomic Scientists, 46(6):46, July/August 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano90-61].

Durrant:1990:RMB

[DW90b] Damian Durrant and Jacqueline Walsh. Report: Military budget: Nuclear weapons for a bygone era. Bulletin of the Atomic Scientists, 46(3):8, 10, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Douglas:1983:CRA

[DWH83] Mary Douglas, Aaron Wildavsky, and John P. Holdren. Commentary: Risk assessment. Bulletin of the Atomic Scientists, 39(7): 58–60, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol83a].

Dycus:1984:L

[Dyc84] J. Stephen Dycus. Law. Bulletin of the Atomic Scientists, 40(10): 17s–18s [VARIOUS between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1959:LEC

[Dys59] Freeman J. Dyson. Letter to the Editor: Caution on ban urged. Bulletin of the Atomic Scientists, 15(10):401, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1960:FDN

[Dys60] Freeman J. Dyson. The future development of nuclear weapons. Foreign Affairs, 38(3):457–464, April 1960. CODEN FR-

NAA3. ISSN 0015-7120. URL http://www.jstor.org/stable/

20029432. Reprinted in [Dys60].

Dyson:1961:RCN

[Dys61] Freeman J. Dyson. Reflections and comments: The neutron bomb. Bulletin of the Atomic Scientists, 17(7):271–272, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1962:TBS

[Dys62] Freeman J. Dyson. Thoughts on bomb shelters. Bulletin of the Atomic Scientists, 18(3):14–15, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Dys63c]. See responses [Mor62a, Ros62].

Dyson:1963:FS

[Dys63a] Freeman Dyson. Forum: On speculation. Bulletin of the Atomic Scientists, 19(7):32, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BB63].

Dyson:1963:DTO

[Dys63b] Freeman J. Dyson. Disarmament through other eyes. Bulletin of the Atomic Scientists, 19(10):37–38, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1963:TBS

[Dys63c] Freeman J. Dyson. Thoughts on bomb shelters. In Grodzins and Rabinowitch [GR63], pages 313–316. LCCN D842 .B78. Reprint of [Dys62].

Dyson:1964:DAB

[Dys64] Freeman J. Dyson. Defense against ballistic missiles. Bulletin of the Atomic Scientists, 20(6):12–18, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [McM65] and reply [Dys65].

Dyson:1965:DR

[Dys65] Freeman J. Dyson. Dyson replies. Bulletin of the Atomic Scientists, 21(3):40, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys64, McM65].

Dyson:1969:CCC

[Dys69a] Freeman J. Dyson. Current comments: A case for missile defense. Bulletin of the Atomic Scientists, 25(4):31–33, April 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ten69, Tay69, You69].

Dyson:1969:HCEa

[Dys69b] Freeman J. Dyson. Human consequences of the exploration of space. Bulletin of the Atomic Scientists, 25(7):8–10, 12–13, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Carlson Memorial Lecture, April 1968. Reprinted in [Dys69c].

Dyson:1969:HCEb

[Dys69c] Freeman J. Dyson. Human consequences of the exploration of space. In Rabinowitch and Lewis [RL69], page ?? LCCN ???? Reprint of [Dys69b].

Dyson:1969:MAM

[Dys69d] Freeman J. Dyson. Missiles and anti-missiles … six views: Comments on Sternglass thesis. Bulletin of the Atomic Scientists, 25 (6):27, June 1969. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://en.wikipedia.org/wiki/
Ernest_J._Sternglass.

Dyson:1972:RBM

[Dys72] Freeman J. Dyson. Review of Ballistic Missile Defense, by Benson D. Adams. Bulletin of the Atomic Scientists, 28(5):41, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1975:HCS

[Dys75] Freeman J. Dyson. The hidden cost of saying no!: Legislators make a mistake in acting as if the future were predictable. Bulletin of the Atomic Scientists, 31(6):23–27, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ehr75, Alb75, Mav75].

Dyson:1981:PCC

[Dys81] Freeman J. Dyson. Perspective: The children’s crusade. Bulletin of the Atomic Scientists, 37(10):1, 4, December 1981. CODEN BASIAP. ISSN 0096-3402. Excerpt from the author’s book, Disturbing the Universe.

Dyson:2007:FI

[Dys07] Freeman Dyson. Freedom of inquiry. Bulletin of the Atomic Scientists, 63(1):62–64, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/62.

Dyson:2020:NB

[Dys20] Freeman Dyson. 1961: The neutron bomb. Bulletin of the Atomic Scientists, 76(6):371–373, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1847491.

Darwin:1947:ARD

[DZA47] Sir Charles Darwin, Ignace Zlotowski, and R. Gordon Arneson. Agency’s research & development activities. Bulletin of the Atomic Scientists, 3(9):256, 266, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

E:1977:MSW

[E.77] S. S. E. Myths spawned within the chemical industry. Bulletin of the Atomic Scientists, 33(3):29–30, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1984:CDW

[E^{+}84] Paul R. Ehrlich et al., editors. The cold and the dark: the world after nuclear war: the Conference on the Long-Term Worldwide Biological Consequences of Nuclear War. Norton, New York, NY, USA, 1984. ISBN 0-393-01870-9. LCCN QH545.N83 C66 1983. Foreword by Lewis Thomas.

Earle:1986:PDA

[Ear86] Ralph Earle II. Perspective: Don’t abandon SALT II. Bulletin of the Atomic Scientists, 42(7):8–9, August/September 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Easwaran:1985:BRB

[Eas85] Eknath Easwaran. Book review: Aborted Discovery: Science and Creativity in the Third World, by Susantha Goonatilake. Bulletin of the Atomic Scientists, 41(5):48, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Easthouse:2000:NCT

[Eas00] Keith Easthouse. Nuclear complex: Too close for comfort. Bulletin of the Atomic Scientists, 56(5):10–12, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/5/10.

Eaton:1962:WG

[Eat62] Anne Eaton. Women in Geneva. Bulletin of the Atomic Scientists, 18(7):47–48, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eaton:1978:DJ

[Eat78] Anne Eaton. Dear Judge. Bulletin of the Atomic Scientists, 34 (4):37–40, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eaves:2016:CCW

[Eav16] Elisabeth Eaves. Crusades of the clueless: Who will win the war on science? Bulletin of the Atomic Scientists, 72(6):418– 421, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eaves:2017:CNA

[Eav17a] Elisabeth Eaves. Can North America’s advanced nuclear reactor companies help save the planet? Bulletin of the Atomic Scientists, 73(1):27–37, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2016.1265353.

Eaves:2017:IDJ

[Eav17b] Elisabeth Eaves. IARPA Director Jason Matheny advances tech tools for US espionage. Bulletin of the Atomic Scientists, 73 (2):67–73, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1288430.

Eaves:2017:NCA

[Eav17c] Elisabeth Eaves. NUKEMAP creator Alex Wellerstein puts nuclear risk on the radar. Bulletin of the Atomic Scientists, 73 (4):211–214, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1338001.

Eaves:2018:GDP

[Eav18] Elisabeth Eaves. Gene drive pioneer Kevin Esvelt tries to make science less secret. Bulletin of the Atomic Scientists, 74(2): 132–138, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1436813.

Eaves:2019:JBR

[Eav19] Elisabeth Eaves. James Brainard: A Republican mayor for city-level climate action. Bulletin of the Atomic Scientists, 75 (2):59–62, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1580872.

English:1973:NGN

[EB73a] Richard D. English and Dan I. Bolef. The numbers game in naval strategic balance. Bulletin of the Atomic Scientists, 29(1):20–25, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pol73] and response [EB73b].

English:1973:CR

[EB73b] Richard D. English and Dan J. Bolef. Communications: Response. Bulletin of the Atomic Scientists, 29(5):4, 48, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [EB73a, Pol73].

Eckholm:1978:SDH

[EB78] Erik Eckholm and Lester R. Brown. Spreading deserts — the hand of man. Bulletin of the Atomic Scientists, 34(1):10–16, 44– 51, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eban:1961:SNS

[Eba61] Abba Eban. Science and the new states. Bulletin of the Atomic Scientists, 17(2):57–60, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eban:1972:NWB

[Eba72] Abba Eban. No way back to Eden. Bulletin of the Atomic Scientists, 28(7):49–51, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

English:1983:CIE

[EBC83] Mary R. English, Robert A. Bohm, and Lillian A. Clinard. Commentary: International Energy Symposia. Bulletin of the Atomic Scientists, 39(4):49, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eberhart:1947:HAP

[Ebe47] Sylvia Eberhart. How the American people feel about the atomic bomb. Bulletin of the Atomic Scientists, 3(6):146–149, 168, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eberhart:2020:HAP

[Ebe20] Sylvia Eberhart. 1947: How the American people feel about the atomic bomb. Bulletin of the Atomic Scientists, 76(6):344– 349, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847472.

Einstein:1952:ESG

[EBU^{+}52] Albert Einstein, Hans A. Bethe, Harold C. Urey, James Franck, Samuel Goudsmit, Cyril S. Smith, Arthur H. Compton, and William P. Murphy. Eminent scientists give their views on American visa policy. Bulletin of the Atomic Scientists, 8(7):217–220, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the Mc-Carran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Eckholm:1978:BBP

[Eck78] Erik Eckholm. Books: Ecoscience: Population, Resources, Environment, by Paul R. Ehrlich, Anne H. Ehrlich, and John P. Holdren. Bulletin of the Atomic Scientists, 34(7):50–51, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Edelson:1958:MOS

[Ede58] Burton I. Edelson. Mutual obligations: Science and the military. Bulletin of the Atomic Scientists, 14(5):169–172, May 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eden:1989:BRB

[Ede89] Lynn Eden. Book review: Minds at War: Nuclear Reality and the Inner Conflicts of Defense Policymakers, by Steven Kull. Bulletin of the Atomic Scientists, 45(3):41–42, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smo89].

Eden:2004:CF

[Ede04] Lynn Eden. City on fire. Bulletin of the Atomic Scientists, 60 (1):32–43, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/1/32.

Eden:2020:CF

[Ede20] Lynn Eden. 2004: City on fire. Bulletin of the Atomic Scientists, 76(6):412–422, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1847516.

Edgerton:1952:QVA

[Edg52] William B. Edgerton. A Quaker view of American foreign policy. Bulletin of the Atomic Scientists, 8(4):111–115, April 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Board:1954:EOC

[Edi54] Editorial Board. Editorial: The Oppenheimer case. Bulletin of the Atomic Scientists, 10(5):173, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

EBB:1960:IJK

[Edi60] Editorial Board of the Bulletin. An interview with John Kennedy. Bulletin of the Atomic Scientists, 16(9):346–347, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Edmonds:1974:TSB

[Edm74] James D. Edmonds, Jr. The tree of science: Beautiful blossoms but bitter fruits. Bulletin of the Atomic Scientists, 30(2):25–26, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dun74].

Edmonds:1975:SES

[Edm75] James D. Edmonds, Jr. Science, engineering and society. Bulletin of the Atomic Scientists, 31(8):34–36, 46, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Edmonds:1977:LS

[Edm77] James D. Edmonds, Jr. Letter: For shame. Bulletin of the Atomic Scientists, 33(2):3, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Edson:1958:AF

[Eds58] James B. Edson. Astronautics and the future. Bulletin of the Atomic Scientists, 14(3):102–105, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [And58].

Edsall:1962:LEW

[Eds62] John T. Edsall. Letter to the Editor: War and politics. Bulletin of the Atomic Scientists, 18(9):30, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Edsall:1976:TPS

[Eds76] John T. Edsall. Toxicity of plutonium and some other actinides. Bulletin of the Atomic Scientists, 32(7):26–37, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and reply [NE76, Coh76b, HE76, Hur77a, Ric77], and correction [Day77d].

Edwards:2012:EHC

[Edw12] Paul N. Edwards. Entangled histories: Climate science and nuclear weapons research. Bulletin of the Atomic Scientists, 68(4): 28–40, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/4/28.

Ehrlich:1981:ESC

[EE81a] Paul Ehrlich and Anne Ehrlich. Extinction or a strategy of conservation. Bulletin of the Atomic Scientists, 37(6):25–30, June/ July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from the authors’ book Extinction: The Causes and Consequences of the Disappearance of Species.

Ehrlich:1981:PE

[EE81b] Paul Ehrlich and Anne Ehrlich. The politics of extinction. Bulletin of the Atomic Scientists, 37(5):26–30, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1985:BRB

[EE85] Paul Ehrlich and Anne Ehrlich. Book review: The Resourceful Earth: A Response to ‘Global 2000’, edited by Julian L. Simon and Herman Kahn. Bulletin of the Atomic Scientists, 41(2):44– 47, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [CEE85].

Ehrlich:1986:PWP

[EE86] Paul R. Ehrlich and Anne H. Ehrlich. Population: World population crisis. Bulletin of the Atomic Scientists, 42(4):13–19, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eger:1978:CPC

[EF78] Martin Eger and Bernard T. Feld. Commentary: The price of collaboration. Bulletin of the Atomic Scientists, 34(10):55–56, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel78f].

Englestein:1986:LSC

[EG86] David Englestein and Carl W. Griffler. Letter: Soviet crisis. Bulletin of the Atomic Scientists, 42(4):53, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cra86a].

Evernden:1989:LGE

[EG89] Jack Evernden and Riley R. Geary. Letter: Not good enough for ban. Bulletin of the Atomic Scientists, 45(5):45–46, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gea89].

Ewing:2024:FTF

[EG24] Rodney C. Ewing and Bernd Grambow. Final thoughts: The fragile connection of safety and science in the geological disposal of radioactive waste. Bulletin of the Atomic Scientists, 81(1):48–52, 2024. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2024.2439761.

Ege:1983:CA

[Ege83] Konrad Ege. Commentary: On ‘Afghanistan’. Bulletin of the Atomic Scientists, 39(10):53–54, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [A N83].

Egeli:2016:SPT

[Ege16] Sitki Egeli. Seeking a path toward missile nonproliferation. Bulletin of the Atomic Scientists, 72(6):362–364, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1972:BDC

[EH72a] Paul R. Ehrlich and John P. Holdren. A Bulletin dialogue: Critique on “The Closing Circle”. Bulletin of the Atomic Scientists, 28(5):16, 18–27, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [Com72, EH72a, EH72b, MW72, But73].

Ehrlich:1972:CEE

[EH72b] Paul R. Ehrlich and John P. Holdren. Communications: The ethics of “Environment”. Bulletin of the Atomic Scientists, 28 (6):6, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehricke:1971:EI

[Ehr71] Krafft A. Ehricke. Extraterrestrial imperative. Bulletin of the Atomic Scientists, 27(9):18–26, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1972:CNQ

[Ehr72] Paul R. Ehrlich. A crying need for quiet conferences: Personal notes from Stockholm. Bulletin of the Atomic Scientists, 28(7): 30–32, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1975:BSY

[Ehr75] Paul R. Ehrlich. The benefits of saying YES! Bulletin of the Atomic Scientists, 31(7):49–51, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys75, Boe76].

Ehrlich:1979:BRB

[Ehr79] Paul R. Ehrlich. Book review: Silent Spring, by Rachel Carson. Bulletin of the Atomic Scientists, 35(8):34–36, October 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1981:BRB

[Ehr81] Anne H. Ehrlich. Book reviews: The Destruction of Nature in the Soviet Union, by Boris Komarov; Overshoot: The Ecological Basis of Revolutionary Change, by William R. Catton, Jr. Bulletin of the Atomic Scientists, 37(9):42–44, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1982:BRBb

[Ehr82a] Anne H. Ehrlich. Book review: International Organization and the Conservation of Nature, by Robert Boardman. Bulletin of the Atomic Scientists, 38(8):45–46, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1982:BRBa

[Ehr82b] Paul R. Ehrlich. Book review: World Population and Human Values: A New Reality, by Jonas Salk and Jonathan Salk. Bulletin of the Atomic Scientists, 38(5):53–54, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1982:DLR

[Ehr82c] Paul R. Ehrlich. Disarmament: the lesser risk. Bulletin of the Atomic Scientists, 38(7):7–8, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1983:BRB

[Ehr83] Anne H. Ehrlich. Book review: Building a Sustainable Society, by Lester R. Brown. Bulletin of the Atomic Scientists, 39(3): 40–42, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ehrlich:1984:NWF

[Ehr84] Anne Ehrlich. Nuclear winter: A forecast of the climatic and biological effects of nuclear war. Bulletin of the Atomic Scientists, 40(4):[various between 24–25], April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eiduson:1966:SAC

[Eid66] Bernice T. Eiduson. Scientists as advisors and consultants in Washington. Bulletin of the Atomic Scientists, 22(8):26–31, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1947:RRM

[Ein47] Albert Einstein. Replies to Ridenour: The military mentality. Bulletin of the Atomic Scientists, 3(8):223–224, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1948:MWC

[Ein48a] Albert Einstein. A message to the World Congress of Intellectuals [Breslau, West Germany, August 1948]. Bulletin of the Atomic Scientists, 4(10):295, 299, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano48-33, Ano48x].

Einstein:1948:PIU

[Ein48b] Albert Einstein. A plea for international understanding. Bulletin of the Atomic Scientists, 4(1):1, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1948:RSS

[Ein48c] Albert Einstein. A reply to the Soviet scientists. Bulletin of the Atomic Scientists, 4(2):35–37, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [VFIS48]. Reprinted in [Ein63].

Einstein:1949:OLF

[Ein49] Albert Einstein. Open letter to friends of the Emergency Committee. Bulletin of the Atomic Scientists, 5(6–7):206, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1950:ACB

[Ein50] Albert Einstein. Arms can bring no security. Bulletin of the Atomic Scientists, 6(3):71, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1951:CM

[Ein51] Albert Einstein. Concern for man. Bulletin of the Atomic Scientists, 7(7–8):204, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1952:MOS

[Ein52a] Albert Einstein. On the moral obligation of the scientist. Bulletin of the Atomic Scientists, 8(2):34–35, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1952:SCD

[Ein52b] Albert Einstein. Symptoms of cultural decay. Bulletin of the Atomic Scientists, 8(7):166–167, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders. Reprinted in [Ein54, pp. 166–167].

Einstein:1953:LRW

[Ein53] Albert Einstein. Letter in reply to William Frauenglass. Bulletin of the Atomic Scientists, 9(6):229–230, July 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Ein54, pages 33–34].

Einstein:1954:IO

[Ein54] Albert Einstein. Ideas and Opinions. Crown Publishers, New York, NY, USA, 1954. 377 pp. LCCN AC35 .E526. Based on Mein Weltbild, edited by Carl Seelig, and other sources. New translations and revisions by Sonja Bargmann.

Einstein:1959:SSA

[Ein59] Albert Einstein. [Schopenhauer, science, and art]. Bulletin of the Atomic Scientists, 15(2):85, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Einstein:1963:ER

[Ein63] Albert Einstein. Einstein replies. In Grodzins and Rabinowitch [GR63], pages 130–134. LCCN D842 .B78. Reply to [VFIS63]. Reprint of [Ein48c].

Einstein:1979:AWW

[Ein79a] Albert Einstein. After World War II. Bulletin of the Atomic Scientists, 35(3):35, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Einstein:1979:MOS

[Ein79b] Albert Einstein. On the moral obligation of the scientist. Bulletin of the Atomic Scientists, 35(3):1, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Eisenhower:1954:CAL

[Eis54a] Dwight D. Eisenhower. Citation: Los Alamos Laboratory. Bulletin of the Atomic Scientists, 10(9):358, November 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1954:PM

[Eis54b] Dwight D. Eisenhower. The President’s message. Bulletin of the Atomic Scientists, 10(4):123–124, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1954:PPA

[Eis54c] Dwight D. Eisenhower. The President’s proposal: Address before the United Nations, December 8, 1953. Bulletin of the Atomic Scientists, 10(1):2–4, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1957:ESS

[Eis57] Dwight D. Eisenhower. Eisenhower speaks on science and security. Bulletin of the Atomic Scientists, 13(10):359, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1958:PS

[Eis58] Dwight D. Eisenhower. The President’s statement. Bulletin of the Atomic Scientists, 14(8):285, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1959:CCN

[Eis59] Dwight D. Eisenhower. Current comment: “Nuclearizing” NATO. Bulletin of the Atomic Scientists, 15(8):347, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1961:EF

[Eis61] Dwight D. Eisenhower. The Eisenhower farewell. Bulletin of the Atomic Scientists, 17(3):90–91, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenhower:1962:FNC

[Eis62] Dwight D. Eisenhower. [The first nuclear chain reaction]. Bulletin of the Atomic Scientists, 18(10):8, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenberg:1966:CHE

[Eis66] Leon Eisenberg. Can human emotions be changed? Bulletin of the Atomic Scientists, 22(1):27–31, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisner:1967:BBN

[Eis67] Robert Eisner. Books: The New Industrial State, by John Kenneth Galbraith. Bulletin of the Atomic Scientists, 23(9):30–32, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisner:1968:WTR

[Eis68] Robert Eisner. War and taxes: The role of the economist in politics. Bulletin of the Atomic Scientists, 24(6):13–18, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisenberg:1972:MRB

[Eis72] Henryk Eisenberg. Macromolecules in research, biology and practice. Bulletin of the Atomic Scientists, 28(8):56–63, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Eisemon:1980:SA

[Eis80] Thomas Owen Eisemon. Scientists in Africa. Bulletin of the Atomic Scientists, 36(2):17–22, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eisner:1984:CDC

[Eis84] Elmer Eisner. Commentary: On ‘The dream of a common security’. Bulletin of the Atomic Scientists, 40(3):64, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei83b].

Eisner:1990:LPA

[Eis90] Elmer Eisner. Letter: Practice arms control techniques on NRA. Bulletin of the Atomic Scientists, 46(7):47, September 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nol90].

Eisendrath:1999:NMI

[Eis99] Craig Eisendrath. Needed: More intelligent intelligence. Bulletin of the Atomic Scientists, 55(6):23–25, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/6/23.

Ekeus:2021:BTW

[Eké21] Rolf Ekéus. Between two wars. Bulletin of the Atomic Scientists, 77(4):163–165, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1941533.

Eklund:1970:IA

[Ekl70] Sigvard Eklund. The international atom. Bulletin of the Atomic Scientists, 26(6):56–61, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eklund:1977:IDG

[Ekl77] Sigvard Eklund. IAEA director-general’s report: ‘We must move forward with all deliberate speed’. Bulletin of the Atomic Scientists, 33(8):42–47, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eklof:1980:BRS

[Ekl80] Ben Eklof. Book reviews: Soviet specialists on America: Soviet Perceptions of the United States, by Morton Schwartz; Russian and the United States, by Nikolaii V. Sivachev and Nikolai N. Yakovlev, translated by Olga Adler Titelbaum. Bulletin of the Atomic Scientists, 36(10):46–51, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eklund:1981:IS

[Ekl81] Sigvard Eklund. The IAEA on safeguards. Bulletin of the Atomic Scientists, 37(8):32–33, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eland:1993:STS

[Ela93] Ivan Eland. Sanctions: Think small. Bulletin of the Atomic Scientists, 49(9):36–40, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elbow:1981:SIB

[Elb81] Richard Ned Elbow. Soviet incentives for brinkmanship? Bulletin of the Atomic Scientists, 37(5):14–21, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eldridge:1992:BRS

[Eld92] Albert F. Eldridge, Jr. Book review: Spymaster’s warning: Terrorism and Democracy, by Stansfield Turner. Bulletin of the Atomic Scientists, 48(3):43–44, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

EPTGKU:1957:CJS

[Ele57] Elementary Particle Theory Group, Kyoto University. Communications from Japanese scientists. Bulletin of the Atomic Scientists, 13(2):73, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eliot:1946:RSD

[Eli46] G. F. Eliot. Russia and the State Department report. Bulletin of the Atomic Scientists, 1(10):8, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Eliassen:1971:PGE

[Eli71] Rolf Eliassen. Power generation and the environment. Bulletin of the Atomic Scientists, 27(7):37–42, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elias:1981:CAB

[Eli81] Richard Elias. Commentary: Atomic bomb 101. Bulletin of the Atomic Scientists, 37(5):63–64, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nor81].

Elias:1985:BRB

[Eli85] Richard Elias. Book review: In a Dark Time, edited by Robert Jay Lifton and Nicholas Humphrey. Bulletin of the Atomic Scientists, 41(6):49–50, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elkinton:1962:SDN

[Elk62] J. Russell Elkinton. Sweet dream or nightmare? Bulletin of the Atomic Scientists, 18(1):34, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elliott:1947:RRF

[Ell47] William Y. Elliott. Replies to Ridenour: Facts and values. Bulletin of the Atomic Scientists, 3(8):227, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elliott:1956:PVS

[Ell56] Walter Elliott. Part V: Science and government: Philosophers and kings. Bulletin of the Atomic Scientists, 12(9):325–328, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ellis:1971:CSU

[Ell71] William N. Ellis. The crisis in science and Unesco. Bulletin of the Atomic Scientists, 27(2):33–35, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ellis:1972:CM

[Ell72] William N. Ellis. Communications: “MINERVA”. Bulletin of the Atomic Scientists, 28(2):48, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LEH^{+}71].

Ellis:1976:LSI

[Ell76] Richard B. Ellis. Letter: Show it all. Bulletin of the Atomic Scientists, 32(2):2, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75g].

Ellis:1977:QR

[Ell77] William N. Ellis. A.T.: the quiet revolution. Bulletin of the Atomic Scientists, 33(9):24–29, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ellsberg:1992:MPI

[Ell92] Daniel Ellsberg. Manhattan Project II. Bulletin of the Atomic Scientists, 48(4):42–44, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elliott:1993:SLR

[Ell93] Kimberly Ann Elliott. Sanctions: A look at the record. Bulletin of the Atomic Scientists, 49(9):32–35, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elmendorf:1965:LP

[Elm65] Armin Elmendorf. Letters: Patents. Bulletin of the Atomic Scientists, 21(2):34, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elmendorf:1966:LEE

[Elm66] Armin Elmendorf. Letter to the Editor: For effective international order. Bulletin of the Atomic Scientists, 22(4):32, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elmendorf:1968:PPW

[Elm68] Armin Elmendorf. Peace and the preparation for war. Bulletin of the Atomic Scientists, 24(8):47–48, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elmendorf:1971:CKD

[Elm71] Armin Elmendorf. Communications: From Kant to Darwin — to 1970. Bulletin of the Atomic Scientists, 27(2):4, 41, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elmendorf:1974:CME

[Elm74] Armin Elmendorf. Communications: In memory of Eugene Rabinowitch. Bulletin of the Atomic Scientists, 30(1):2, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elmendorf:1976:LPP

[Elm76] Armin Elmendorf. Letter: Pied pipers. Bulletin of the Atomic Scientists, 32(2):3, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elshtain:1990:BRT

[Els90] Jean Bethke Elshtain. Book review: Think again: Unthinking the Unthinkable: Nuclear Weapons and Western Culture, by Jeff Smith. Bulletin of the Atomic Scientists, 46(10):48, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Elwyn:1964:LED

[Elw64] David H. Elwyn. Letter to the Editor: DNA. Bulletin of the Atomic Scientists, 20(8):36–37, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul64a].

Emanuel:2021:NFI

[Ema21] Kerry Emanuel. Nuclear fear: The irrational obstacle to real climate action. Bulletin of the Atomic Scientists, 77(6):285–289, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1989192.

ECAS:1947:SRD

[Eme47] Emergency Committee of Atomic Scientists. Scientists re-define educational aims. Bulletin of the Atomic Scientists, 3(8):216– 217, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ECAS:1948:PSS

[Eme48] Emergency Committee of Atomic Scientists. A policy for survival: A statement by Emergency Committee of Atomic Scientists — April 12, 1948. Bulletin of the Atomic Scientists, 4(6):176, 188, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emery:1954:LER

[Eme54] Alden H. Emery. Letter to the Editor: Response to M. Haissinsky. Bulletin of the Atomic Scientists, 10(2):56, February 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hai54] and comment [Tod55].

Emelyanov:1960:TCI

[Eme60] V. S. Emelyanov. Toward close international cooperation in atomic research. Bulletin of the Atomic Scientists, 16(8):322– 325, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emelyanov:1961:NSN

[Eme61] V. Emelyanov. The neutron should not be neutral. Bulletin of the Atomic Scientists, 17(9):389–390, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emelyanov:1963:APD

[Eme63] V. Emelyanov. Atomic power and disarmament. Bulletin of the Atomic Scientists, 19(8):16–20, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emelyanov:1965:MPW

[Eme65] Vasily S. Emelyanov. For more pure water. Bulletin of the Atomic Scientists, 21(7):46–48, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emelyanov:1971:NES

[Eme71a] V. S. Emelyanov. Nuclear energy in the Soviet Union. Bulletin of the Atomic Scientists, 27(9):37–41, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emerson:1971:OEU

[Eme71b] John D. Emerson. Outlook for energy in the United States. Bulletin of the Atomic Scientists, 27(8):18–19, October 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emelyanov:1976:NPP

[Eme76] V. S. Emelyanov. Nuclear power in peace and war. Bulletin of the Atomic Scientists, 32(3):18–19, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Emlen:1966:BBC

[Eml66] John T. Emlen. Books: A Continent for Science — The Antarctic Adventure, by Richard S. Lewis. Bulletin of the Atomic Scientists, 22(2):37, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Endicott:2000:IR

[End00] John E. Endicott. Intellectual rigor. Bulletin of the Atomic Scientists, 56(2):73–74, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/2/73.

Engelke:1978:CSC

[Eng78] Charles E. Engelke. Commentary: A self-contained community energy system. Bulletin of the Atomic Scientists, 34(9):51–53, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [IES79].

English:1988:BRB

[Eng88] Robert D. English. Book review: Nuclear Blackmail and Nuclear Balance, by Richard K. Betts. Bulletin of the Atomic Scientists, 44(6):52, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CE88].

Engel:1991:BRD

[Eng91] Marguerite Engel. Book review: Day’s long journey: Crossing the Line, by Samuel H. Day, Jr. Bulletin of the Atomic Scientists, 47(8):44–45, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Engelman:2008:PSP

[Eng08] Robert Engelman. A population strategy for the planet. Bulletin of the Atomic Scientists, 64(5):44, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/5/44.

Evstafiev:2000:RPP

[EO00] Dimitry G. Evstafiev and Vladimir A. Orlov. Russia’s political players. Bulletin of the Atomic Scientists, 56(2):57–62, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/2/57.

Eslami:1986:LCN

[EP86] Al S. Eslami and Tony Prudori. Letter: Conditional no first use? Bulletin of the Atomic Scientists, 42(4):55–56, April 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Plo86].

Eppich:2020:NFH

[Epp20] Gary R. Eppich. Nuclear forensics: How science helps stop the trafficking of nuclear materials. Bulletin of the Atomic Scientists, 76(5):263–270, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1806587.

Epstein:1952:VDM

[Eps52] Paul S. Epstein. VIII: The diamat and modern science. Bulletin of the Atomic Scientists, 8(6):190–194, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See opposing view in [Hol53].

Epstein:1975:LFU

[Eps75a] Kenneth J. Epstein. Letters: Fear of the unknown. Bulletin of the Atomic Scientists, 31(8):4, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1975:FNR

[Eps75b] William Epstein. Failure at NPT Review Conference. Bulletin of the Atomic Scientists, 31(7):46–48, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1975:IRM

[Eps75c] William Epstein. The inexorable rise of military expenditures: Despite detente, the budgets always go up. Bulletin of the Atomic Scientists, 31(1):17–19, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1977:LBA

[Eps77a] Kenneth J. Epstein. Letter: A better arms race. Bulletin of the Atomic Scientists, 33(9):2, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1977:CE

[Eps77b] Samuel S. Epstein. Cancer and the environment. Bulletin of the Atomic Scientists, 33(3):22–28, 30, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1977:CCP

[Eps77c] Samuel S. Epstein. The case for a consumer protection agency. Bulletin of the Atomic Scientists, 33(7):6–7, September 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1977:TBD

[Eps77d] William Epstein. Time to bury ‘deterrence’. Bulletin of the Atomic Scientists, 33(6):6–7, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1979:CBU

[Eps79] William Epstein. Commentary: Banning the use of nuclear weapons. Bulletin of the Atomic Scientists, 35(4):7–9, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1981:CSR

[Eps81] William Epstein. Commentary: On second review of Non-Proliferation Treaty. Bulletin of the Atomic Scientists, 37(5): 57–60, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1982:FHI

[Eps82a] William Epstein. The freeze: a hot issue at the United Nations. Bulletin of the Atomic Scientists, 38(8):50–52, October 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1982:UNI

[Eps82b] William Epstein. The United Nations and international security. Bulletin of the Atomic Scientists, 38(6):37–40, June 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bon83].

Epstein:1986:NHC

[Eps86] William Epstein. New hope for a comprehensive test ban. Bulletin of the Atomic Scientists, 42(2):29–30, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1987:CDN

[Eps87a] William Epstein. Canadian defense: New stance tarnishes Canada’s reputation. Bulletin of the Atomic Scientists, 43(8): 11–12, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1987:PNT

[Eps87b] William Epstein. Perspective: Nuclear testing: illusion and reality. Bulletin of the Atomic Scientists, 43(3):8, April 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1988:LPQ

[Eps88a] Kenneth J. Epstein. Letter: Protest quantitatively. Bulletin of the Atomic Scientists, 44(7):62, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coa88].

Epstein:1988:PUP

[Eps88b] William Epstein. Perspective: U.N. presses superpowers on test ban. Bulletin of the Atomic Scientists, 44(2):7–8, March 1988.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1990:BTS

[Eps90a] William Epstein. Bulletin: Thou shalt ban testing. Bulletin of the Atomic Scientists, 46(2):5–6, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1990:NRC

[Eps90b] William Epstein. NPT review: Conference a qualified success. Bulletin of the Atomic Scientists, 46(10):45–47, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1990:NRN

[Eps90c] William Epstein. NPT review: The nuclear testing threat. Bulletin of the Atomic Scientists, 46(6):34–37, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1991:RTB

[Eps91] William Epstein. Report: Test ban: January meeting keeps hope alive. Bulletin of the Atomic Scientists, 47(3):10–11, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1992:NUC

[Eps92a] William Epstein. And now — the U.N. century. Bulletin of the Atomic Scientists, 48(4):22–23, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1992:RUN

[Eps92b] William Epstein. Report: United Nations: Write down your arms. Bulletin of the Atomic Scientists, 48(2):11–12, 44, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1993:BCT

[Eps93] William Epstein. Bulletin: CTB: Two paths, one goal. Bulletin of the Atomic Scientists, 49(8):3–5, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1994:SCG

[Eps94] William Epstein. Stand and be counted: Give more to get more. Bulletin of the Atomic Scientists, 50(6):15–18, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Wil95, DE95].

Epstein:1995:NWI

[Eps95] William Epstein. NPT wrap-up: Indefinite extension — with increased accountability. Bulletin of the Atomic Scientists, 51(4): 27–30, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1996:CNS

[Eps96a] William Epstein. CTB: Next steps. Bulletin of the Atomic Scientists, 52(6):36–37, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1996:RNA

[Eps96b] William Epstein. Revolt of the non-aligned. Bulletin of the Atomic Scientists, 52(2):39–40, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1997:RUN

[Eps97] William Epstein. Report: United Nations: Nuclear powers take a hard line. Bulletin of the Atomic Scientists, 53(2):13–15, March/ April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:2012:BYC

[Eps12] Gerald L. Epstein. Biosecurity 2011: Not a year to change minds. Bulletin of the Atomic Scientists, 68(1):29–38, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/1/29.

Epstein:1984:UDC

[ER84] William Epstein and Alfonso Garcia Robles. U.N. disarmament campaign. Bulletin of the Atomic Scientists, 40(3):37–39, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ebel:1990:LSS

[ER90] Paul Ebel and Martin Rees. Letters: Satellites’ siren song. Bulletin of the Atomic Scientists, 46(1):47, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [JR89].

Erdos:1952:SBE

[Erd52] Paul Erdős. Some British experiences. Bulletin of the Atomic Scientists, 8(7):231, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Ericson:1988:SNGa

[Eri88a] R. E. Ericson. Soviet numbers game threatens perestroika. Bulletin of the Atomic Scientists, 44(10):20, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ericson:1988:SNGb

[Eri88b] Richard E. Ericson. Soviet numbers game threatens perestroika. Bulletin of the Atomic Scientists, 44(10):20–25, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Erikson:1991:RLD

[Eri91] Kai Erikson. Radiation’s lingering dread. Bulletin of the Atomic Scientists, 47(2):34–39, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ericson:1995:BRW

[Eri95] Richard E. Ericson. Book review: Welfare state: The Soviet Social Contract and Why It Failed, by Linda J. Cook. Bulletin of the Atomic Scientists, 51(2):67–68, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Errera:1958:WDU

[Err58] Maurice Errera. What did the U.N. Radiation Committee accomplish? Bulletin of the Atomic Scientists, 14(9):388–393, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Epstein:1989:NNS

[ES89] William Epstein and Glenn T. Seaborg. Non-nuclear states move to end testing. Bulletin of the Atomic Scientists, 45(5):36–37, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Esposito:1972:PMP

[Esp72a] Bruce J. Esposito. The politics of medicine in the People’s Republic of China. Bulletin of the Atomic Scientists, 28(10):4–9, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Esposito:1972:SMC

[Esp72b] Bruce J. Esposito. Science in Mainland China. Bulletin of the Atomic Scientists, 28(1):36–40, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rot72].

Espona:2014:WBA

[Esp14] Maria José Espona. Winning the battle against emerging pathogens: an Argentine response. Bulletin of the Atomic Scientists, 70(4):18–21, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Estrella:2001:AMS

[Est01] Hugo Estrella. Argentina: Menem’s star falls. Bulletin of the Atomic Scientists, 57(5):13–15, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/13.

Einstein:2020:WSS

[ET20] Albert Einstein and Edward Teller. 1950: What the scientists are saying. Bulletin of the Atomic Scientists, 76(6):350–352, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847477.

Etzioni:1963:FGA

[Etz63] Amitai Etzioni. Forum: Gradual arms reduction. Bulletin of the Atomic Scientists, 19(8):30–33, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Etzioni:1964:FWS

[Etz64] Amitai Etzioni. Forum: When scientists testify. Bulletin of the Atomic Scientists, 20(8):23–26, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Etzioni:1977:CRO

[Etz77] Amitai Etzioni. A creative response to our crisis. Bulletin of the Atomic Scientists, 33(2):24, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Etzioni:2007:LPU

[Etz07] Amitai Etzioni. A less perfect union. Bulletin of the Atomic Scientists, 63(3):19–21, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/3/19.

Edsall:1984:CHR

[EV84] John T. Edsall and Theodore H. Von Laue. Commentary: On ’Human rights imperialism’. Bulletin of the Atomic Scientists, 40(1):58–60, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Von83].

Evans:1954:MRO

[Eva54] Ward V. Evans. Minority report [on Oppenheimer security clearance denial]. Bulletin of the Atomic Scientists, 10(6):250, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evans:1963:WNP

[Eva63] Gordon Heyd Evans. The world of nuclear plenty. Bulletin of the Atomic Scientists, 19(2):26–30, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evans:1975:DVP

[Eva75] Gary W. Evans. A dialogue on violence: Pacificism does not mean passivity. Bulletin of the Atomic Scientists, 31(10):46–47, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Fictional dialogue between Frantz Fanon and Mohandas Gandhi.

Evangelista:1982:CHR

[Eva82] Matthew A. Evangelista. Commentary: Human rights linkage. Bulletin of the Atomic Scientists, 38(3):47–49, March 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor81, CR81].

Evan:1984:COH

[Eva84] William M. Evan. Commentary: One-half of one percent for peace. Bulletin of the Atomic Scientists, 40(4):39, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evangelista:1986:BRB

[Eva86] Matthew Evangelista. Book review: The Soviet Calculus of Nuclear War, edited by Roman Kolkowicz and Ellen Propper Mickiewicz. Bulletin of the Atomic Scientists, 42(10):47–48, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evangelista:1987:EST

[Eva87a] Matthew Evangelista. Exploiting the Soviet “threat” to Europe. Bulletin of the Atomic Scientists, 43(1):14–16, 18, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evangelista:1987:SG

[Eva87b] Matthew Evangelista. Sakharov and Gorky. Bulletin of the Atomic Scientists, 43(3):21, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evangelista:1999:RFU

[Eva99] Matthew Evangelista. Russia’s fragile union. Bulletin of the Atomic Scientists, 55(3):50–55, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/3/50.

Evans:2004:MDW

[Eva04] Nicole C. Evans. Missile defense: Winning minds, not hearts. Bulletin of the Atomic Scientists, 60(5):48–55, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/5/48.

Evernden:1985:PTT

[Eve85] Jack Evernden. Politics, technology, and the test ban. Bulletin of the Atomic Scientists, 41(3):9–12, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evernden:1988:LST

[Eve88a] Jack F. Evernden. Lies that stopped a test ban. Bulletin of the Atomic Scientists, 44(8):20–24, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bro88b, Eve88b, PE89].

Evernden:1988:RTT

[Eve88b] Jack F. Evernden. Round two of test ban flap: Yes you did. Bulletin of the Atomic Scientists, 44(10):6, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eve88a, Bro88b].

Everts:1989:WPM

[Eve89] Philip P. Everts. Where the peace movement goes when it disappears. Bulletin of the Atomic Scientists, 45(9):26–30, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Evron:1978:ASM

[Evr78] Yair Evron. Arms and security in the Middle East. Bulletin of the Atomic Scientists, 34(2):44–49, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Edmonds:2006:TT

[EW06] Sibel Edmonds and William Weaver. To tell the truth. Bulletin of the Atomic Scientists, 62(1):72, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/72.

Ewing:1960:LEP

[Ewi60] Shirley Ewing. Letter to the Editor: A plaudit for the Bulletin. Bulletin of the Atomic Scientists, 16(1):33, January 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BW59].

ECFAS:1954:SDO

[Exe54] Executive Committee of the Federation of American Scientists. The suspension of Dr. Oppenheimer. Bulletin of the Atomic Scientists, 10(5):191–192, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ECFCNL:1957:CDP

[Exe57] Executive Council of the Friends of Committee on National Legislation. Civil defense and peace: A Quaker view. Bulletin of the Atomic Scientists, 13(5):176, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ezeh:2016:PPM

[Eze16] Alex Ezeh. Population’s part in mitigating climate change: A Nigerian response. Bulletin of the Atomic Scientists, 72(3):189– 191, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ezrahi:1972:SC

[Ezr72] Yaron Ezrahi. The slaves of the 20th Century. Bulletin of the Atomic Scientists, 28(10):10–11, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jan73].

F:1961:CCN

[F.61] P. F. Current comment: New peace corps. Bulletin of the Atomic Scientists, 17(4):160–161, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fahmy:1990:PED

[Fah90] Mohamed Nabil Fahmy. Perspective: Egypt’s disarmament initiative. Bulletin of the Atomic Scientists, 46(9):9–10, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fahmy:1991:NWO

[Fah91] Mohamed Nabil Fahmy. New world order: Obey the U.N. Charter. Bulletin of the Atomic Scientists, 47(5):30–31, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fainberg:1979:CMT

[Fai79] Anthony Fainberg. Commentary: More on Three Mile Island. Bulletin of the Atomic Scientists, 35(9):65, November 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Nus80].

Fainberg:1980:GRP

[Fai80] Anthony Fainberg. Ground rules for the power debate. Bulletin of the Atomic Scientists, 36(3):44–46, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Arn80].

Fainberg:1981:OIS

[Fai81] Anthony Fainberg. Osirak and international security. Bulletin of the Atomic Scientists, 37(8):33–36, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fairbank:1983:F

[Fai83] John K. Fairbank. Furthermore. Bulletin of the Atomic Scientists, 39(2):15, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Orl83].

Fai:1997:LUC

[Fai97] Ghulam Nabi Fai. Letter: U.S. can help end Kashmir conflict. Bulletin of the Atomic Scientists, 53(3):61, May/June 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Obe97a].

Fallers:1961:CPS

[Fal61a] Lloyd Fallers. C. P. Snow and the third culture. Bulletin of the Atomic Scientists, 17(8):306–310, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [CMMF62].

Fallers:1961:BBN

[Fal61b] Lloyd A. Fallers. Books: The New Nigerian Elite, by Hugh H. Smythe and Mabel M. Smythe. Bulletin of the Atomic Scientists, 17(2):76, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Falk:1968:SPP

[Fal68] Charles E. Falk. Science and public policy activities in universities. Bulletin of the Atomic Scientists, 24(6):50–52, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Falicov:1970:PPL

[Fal70] L. M. Falicov. Physics and politics in Latin America — a personal experience. Bulletin of the Atomic Scientists, 26(9):8–10, 41–45, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bon72].

Falk:1979:PHN

[Fal79] Richard A. Falk. Panel I: How a nuclear war can start … in the Middle East. Bulletin of the Atomic Scientists, 35(4):21–22, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Falk:1985:LML

[Fal85] Richard Falk. Liberation from military logic. Bulletin of the Atomic Scientists, 41(7):136–139, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Falk:1987:IVO

[Fal87] Richard Falk. International views: An opportunity to end the arms race. Bulletin of the Atomic Scientists, 43(2):28–32, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pes87].

Falk:1989:ILD

[Fal89] Richard Falk. International law: The damaged U.S. image. Bulletin of the Atomic Scientists, 45(1):59–61, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Farmer:1971:BVS

[Far71] F. R. Farmer. A British view: Safety and nuclear power plants. Bulletin of the Atomic Scientists, 27(9):47–49, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fartash:1977:DCW

[Far77] M. Fartash. The ‘disarmament club’ at work. Bulletin of the Atomic Scientists, 33(1):57–62, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Farrell:1986:LBB

[Far86] James J. Farrell. Letter: By the Bomb’s Early Light. Bulletin of the Atomic Scientists, 42(7):65, August/September 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [For86].

Farley:1987:HNT

[Far87] Philip J. Farley. How to negotiate a treaty. Bulletin of the Atomic Scientists, 43(8):33–36, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Farinella:1998:LCN

[Far98] Paolo Farinella. Letter: On a collision with near-Earth reality. Bulletin of the Atomic Scientists, 54(6):3, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot98e].

Farseta:2008:CSN

[Far08] Diane Farseta. The campaign to sell nuclear. Bulletin of the Atomic Scientists, 64(4):??, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/4/38.

Faro:2013:LLN

[Far13] Alexandre Faro. The legalities of leaving nuclear. Bulletin of the Atomic Scientists, 69(1):36–42, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/1/36.

Fasteau:1966:FMV

[Fas66] Marc S. Fasteau. Forum: Munich and Vietnam: A valid analogy? Bulletin of the Atomic Scientists, 22(7):22–25, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fascell:1978:CBS

[Fas78] Dante B. Fascell. Commentary: Breaking the silence barrier. Bulletin of the Atomic Scientists, 34(8):5–7, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Siemes:1946:HA

[Fat46] Father Siemes, S.J. Hiroshima — August 6, 1945. Bulletin of the Atomic Scientists, 1(11):2–6, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Faulk:1987:LRS

[Fau87] Stuart Faulk. Letter: Real Soviet fears. Bulletin of the Atomic Scientists, 43(3):53, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro86e].

Favre:1952:NSP

[Fav52] Christian Favre. Neuro-surgery and the problems it raises. Bulletin of the Atomic Scientists, 8(6):178, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fay:1975:BBB

[Fay75] James A. Fay. Books: The Brain Bank of America: An Inquiry Into the Politics of Science, by Philip Boffey. Bulletin of the Atomic Scientists, 31(7):55–56, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1989:CSC

[FB89] David Fischer and Richard Bolt. Counterpoint: Safeguards controversy, continued. Bulletin of the Atomic Scientists, 45(5):38– 40, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bol88a].

Fainberg:1981:CLU

[FBH81] Anthony Fainberg, Alan M. Bieber, and William A. Higinbotham. Commentary: On ‘Life under the gun’. Bulletin of the Atomic Scientists, 37(5):62–63, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alb81].

Fricaud-Chagnaud:1984:FD

[FC84] Georges Fricaud-Chagnaud. The French defense. Bulletin of the Atomic Scientists, 40(5):14–15, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kol84].

Fricaud-Chagnaud:1986:FNU

[FC86] George Fricaud-Chagnaud. France’s nuclear umbrella. Bulletin of the Atomic Scientists, 42(5):33–36, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fricaud-Chagnaud:1988:ASF

[FC88] Georges Fricaud-Chagnaud. After the summit: The French foundation for European security. Bulletin of the Atomic Scientists, 44(2):24–25, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fricaud-Chagnaud:1996:LEN

[FC96] C. G. Fricaud-Chagnaud. Letter: Eurobomb? Non; Eurodeterrence? Oui. Bulletin of the Atomic Scientists, 52(3):3, 60, May/ June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hib96b] and comment [HW96].

Finch:1946:CDC

[FCC^{+}46] George A. Finch, Joseph P. Chamberlain, Percy E. Corbett, Malcolm W. Davis, Clyde Eagleton, Manley O. Hudson, Herbert L. May, James T. Shotwell, Edgar Turlington, and Louis B. Sohn. The Carnegie draft convention … legal subcommittee, Carnegie Endowment for International Peace. Bulletin of the Atomic Scientists, 2(1–2):15–19, July 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fulton:1976:LDA

[FD76] Christopher B. Fulton and Samuel H. Day, Jr. Letter: Destructive attitudes. Bulletin of the Atomic Scientists, 32(5):2, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ben76b].

Fingar:1978:SPP

[FD78] Thomas Fingar and Genevieve Dean. Science policy in the PRC. Bulletin of the Atomic Scientists, 34(8):60–61, October 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Funabashi:2023:FLL

[FD23] Yoichi Funabashi and Marina Fujita Dickson. Fukushima: Lessons learned from a devastating “near-miss”. Bulletin of the Atomic Scientists, 79(3):161–165, 2023. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2200121.

Fisk:1952:CMP

[FDC52] J. B. Fisk, Farrington Daniels, and J. D. Cockcroft. Comments on Mr. Pitzer’s speech. Bulletin of the Atomic Scientists, 8(8): 288, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pit52b].

Fraise:2023:AAH

[FE23] Thomas Fraise and Kjølv Egeland. Able archer: How close of a call was it? Bulletin of the Atomic Scientists, 79(3):155–160, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2199577.

FAS:1950:BD

[Fed50] Federation of Atomic Scientists. Break the deadlock! Bulletin of the Atomic Scientists, 6(3):74, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1955:AVP

[Fed55] Federation of Atomic Scientists. American visa policy: A report. Bulletin of the Atomic Scientists, 11(10):367–368, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1958:NTB

[Fed58] Federation of American Scientists. Nuclear test ban, U.N. control of space research, and U.N. police force — first steps toward peace. Bulletin of the Atomic Scientists, 14(3):125, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fedorov:1959:ACN

[Fed59a] E. K. Fedorov. The agreement on the cessation of nuclear tests must be concluded without delay! Bulletin of the Atomic Scientists, 15(8):329–332, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fedorov:1959:TPCa

[Fed59b] E. K. Fedorov. Third Pugwash Conference: Controlled cessation of atomic weapons tests. Bulletin of the Atomic Scientists, 15(1): 8–11, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fedorov:1959:TPCb

[Fed59c] E. K. Fedorov. Third Pugwash Conference: Some lessons of the International Geophysical Year. Bulletin of the Atomic Scientists, 15(4):155–159, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FASUNUC:1961:RCN

[Fed61] Federation of American Scientists U.N.University Committee. Readers’ comments: The need for a U.N. university. Bulletin of the Atomic Scientists, 17(3):111–113, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1962:CDS

[Fed62] Federation of American Scientists. Civil defense shelter statement. Bulletin of the Atomic Scientists, 18(2):25–28, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1963:FFA

[Fed63] Federation of American Scientists. FAS: For further arms control. Bulletin of the Atomic Scientists, 19(9):46, November 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1964:FSB

[Fed64] Federation of American Scientists. FAS statement on biological and chemical warfare. Bulletin of the Atomic Scientists, 20(8): 46–47, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1967:CWW

[Fed67a] Federation of American Scientists. Comments: The war and weapons in Vietnam. Bulletin of the Atomic Scientists, 23(5): 59–60, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1967:RCR

[Fed67b] Federation of American Scientists. Reports: Classified research in the university. Bulletin of the Atomic Scientists, 23(8):45–46, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1968:SCB

[Fed68a] Federation of American Scientists. Scientists on C and B. Bulletin of the Atomic Scientists, 24(7):31–32, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1968:SS

[Fed68b] Federation of American Scientists. Scientists speak out. Bulletin of the Atomic Scientists, 24(5):50, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fedorov:1970:AEP

[Fed70] Ye. K. Fedorov. Antarctica: Experimental proving ground for peaceful coexistence and international collaboration. Bulletin of the Atomic Scientists, 26(10):22–28, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Fedorov:1972:IME

[Fed72] E. K. Fedorov. The interaction of man and the environment: The situation today and the outlook for the future. Bulletin of the Atomic Scientists, 28(2):5–10, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

FAS:1976:FUD

[Fed76] Federation of American Scientists. First use deserves more than one decision-maker. Bulletin of the Atomic Scientists, 32(3):54– 55, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Federow:1978:RDN

[Fed78] Harold L. Federow. Recombinant DNA and nuclear energy. Bulletin of the Atomic Scientists, 34(2):6–7, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feinstein:1960:ASA

[Fei60] Otto Feinstein. American scholars analyze U.S. foreign policy. Bulletin of the Atomic Scientists, 16(10):395–399, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feinstein:1961:ASA

[Fei61] Otto Feinstein. American scholars analyze U.S. foreign policy. Bulletin of the Atomic Scientists, 17(1):30–34, January 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feinberg:1971:SYW

[Fei71] Gerald Feinberg. Survival? Yes. But in what form? Bulletin of the Atomic Scientists, 27(5):27–30, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feiveson:1977:BBD

[Fei77] Harold A. Feiveson. Books: The breeder debate Nuclear Reactors: To Breed or Not to Breed, edited by J. Rotblat. Bulletin of the Atomic Scientists, 33(6):63–64, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feierman:1981:SSC

[Fei81a] Barry H. Feierman. Survival of the species: a classroom project. Bulletin of the Atomic Scientists, 37(9):20–23, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fein:1981:CSW

[Fei81b] Elihu Fein. Commentary: The sacred weapons. Bulletin of the Atomic Scientists, 37(7):52–53, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feiveson:1981:CFM

[Fei81c] H. A. Feiveson. Commentary: Further more …. Bulletin of the Atomic Scientists, 37(7):10–11, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ne’81].

Feinstein:1990:CWS

[Fei90] Lee Feinstein. Chemical weapons: Speak loudly, carry a small stick. Bulletin of the Atomic Scientists, 46(10):8, 48–49, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feinstein:1992:RAC

[Fei92] Lee Feinstein. Report: Arms control: Arms R Us. Bulletin of the Atomic Scientists, 48(9):8–10, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feiveson:2007:WE

[Fei07] H. A. Feiveson. A wasteful endeavor. Bulletin of the Atomic Scientists, 63(5):57–59, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/5/57.

Feld:1955:LAC

[Fel55] Bernard T. Feld. Let’s abolish classification in the atomic power field. Bulletin of the Atomic Scientists, 11(6):219–220, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fellows:1957:LEL

[Fel57] Margaret A. Fellows. Letter to the Editor: A long hard look at civil defense. Bulletin of the Atomic Scientists, 13(1):41, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim56].

Feld:1961:RCD

[Fel61] Bernard T. Feld. Readers’ comment: Disarmament or arms control — a matter of semantics. Bulletin of the Atomic Scientists, 17(5–6):244, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1962:ECA

[Fel62a] Bernard T. Feld. Events and comments: Atmospheric testing. Bulletin of the Atomic Scientists, 18(1):31–33, January 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1962:LEA

[Fel62b] Bernard T. Feld. Letter to the Editor: Atmospheric testing. Bulletin of the Atomic Scientists, 18(6):43–44, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rat62].

Feld:1962:MIT

[Fel62c] Bernard T. Feld. More important than shelters. Bulletin of the Atomic Scientists, 18(4):8–11, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1964:NNW

[Fel64] Bernard T. Feld. The nonproliferation of nuclear weapons. Bulletin of the Atomic Scientists, 20(10):2–5, December 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1966:AN

[Fel66a] Bernard T. Feld. Author’s note. Bulletin of the Atomic Scientists, 22(5):32, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1966:BNB

[Fel66b] Bernard T. Feld. Books: The Nagasaki binge: Day of Trinity, by Lansing Lamont. Bulletin of the Atomic Scientists, 22(2):35– 37, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See apology [Fel66a].

Feld:1966:LEC

[Fel66c] Bernard T. Feld. Letter to the Editor: On the Chinese separation technology. Bulletin of the Atomic Scientists, 22(7):33–34, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Felix:1966:LEP

[Fel66d] David Felix. Letter to the Editor: Privilege and poverty. Bulletin of the Atomic Scientists, 22(10):29–30, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1967:CNT

[Fel67a] B. T. Feld. Comments: The Nonproliferation Treaty: An acceptable balance. Bulletin of the Atomic Scientists, 23(9):37–38, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1967:DDE

[Fel67b] B. T. Feld. The decision to deploy: An editorial opinion. Bulletin of the Atomic Scientists, 23(10):18–19, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WF68, BJH68].

Feld:1967:CN

[Fel67c] Bernard T. Feld. Comments: To be nuclear or not. Bulletin of the Atomic Scientists, 23(5):60–62, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1967:PNF

[Fel67d] Bernard T. Feld. A pledge: No first use. Bulletin of the Atomic Scientists, 23(5):46–48, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1968:ANT

[Fel68a] Bernard T. Feld. After the Nonproliferation Treaty — what next? Bulletin of the Atomic Scientists, 24(7):2–3, September 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1968:BBP

[Fel68b] Bernard T. Feld. Books: Unless Peace Comes, edited by Nigel Calder. Bulletin of the Atomic Scientists, 24(9):35–36, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Law69, Fel69a].

Feld:1968:CCN

[Fel68c] Bernard T. Feld. Current comments: The Nuclear Nonproliferation Treaty — in the cards for 1968? Bulletin of the Atomic Scientists, 24(3):34–35, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1968:TNA

[Fel68d] Bernard T. Feld. Toward a new American program for peace. Bulletin of the Atomic Scientists, 24(5):4–5, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Nat68a].

Feld:1969:FR

[Fel69a] Bernard T. Feld. Feld responds. Bulletin of the Atomic Scientists, 25(3):31, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel68b, Law69].

Feld:1969:ORS

[Fel69b] Bernard T. Feld. Other reactions in the science community. Bulletin of the Atomic Scientists, 25(6):48, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1970:SRA

[Fel70a] Bernard T. Feld. Scientist’s role in arms control. Bulletin of the Atomic Scientists, 26(1):7–8, 47–48, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1970:SHA

[Fel70b] Bernard T. Feld. The sorry history of arms control. Bulletin of the Atomic Scientists, 26(7):22–26, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1971:ECB

[Fel71] Bernard T. Feld. An editorial: China and the bomb. Bulletin of the Atomic Scientists, 27(7):0, 31, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1972:IAL

[Fel72a] Bernard T. Feld. Issues in arms limitations — four views: III. Current developments and dangers of atomic armaments. Bulletin of the Atomic Scientists, 28(3):16–20, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1972:LSI

[Fel72b] Bernard T. Feld. Looking to SALT-II. Bulletin of the Atomic Scientists, 28(6):2–3, 50, 54, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1973:BBT

[Fel73a] Bernard T. Feld. Books: Science and Technology as an Instrument of Soviet Policy, by Mose L. Harvey, Leon Goure, and V. Prokofieff. Bulletin of the Atomic Scientists, 29(9):52–54, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1973:LAM

[Fel73b] Bernard T. Feld. Lev A. Artsimovich: In memoriam. Bulletin of the Atomic Scientists, 29(4):43, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1973:TI

[Fel73c] Bernard T. Feld. A ticket to Istanbul. Bulletin of the Atomic Scientists, 29(6):5, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1973:VCS

[Fel73d] Bernard T. Feld. A voice of conscience is stilled. II. origins of Pugwash. Bulletin of the Atomic Scientists, 29(6):4–6, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1974:GEM

[Fel74a] Bernard T. Feld. Guest editorial: More energy at what price? Bulletin of the Atomic Scientists, 30(5):39–41, May 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1974:TAN

[Fel74b] Bernard T. Feld. Three aspects of the nuclear problem: The menace of a fission power economy. Bulletin of the Atomic Scientists, 30(4):32–34, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hur74].

Feld:1975:ACW

[Fel75a] Bernard T. Feld. Arms control or what? Bulletin of the Atomic Scientists, 31(9):5–7, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wal76].

Feld:1975:CPA

[Fel75b] Bernard T. Feld. The charade of piecemeal arms limitation: Time has come for doves of the world to unite behind a comprehensive agreement. Bulletin of the Atomic Scientists, 31(1):8–16, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1975:EMW

[Fel75c] Bernard T. Feld. Editorial: Making the world safe for plutonium. Bulletin of the Atomic Scientists, 31(5):5–6, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:CNW

[Fel76a] Bernard T. Feld. Consequences of nuclear war. Bulletin of the Atomic Scientists, 32(6):10–13, June 1976. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lor77].

Feld:1976:ENB

[Fel76b] Bernard T. Feld. Editorial: A new ball game? Bulletin of the Atomic Scientists, 32(3):6–7, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:ECP

[Fel76c] Bernard T. Feld. Editorial: Congress passes the buck. Bulletin of the Atomic Scientists, 32(8):5, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:EGS

[Fel76d] Bernard T. Feld. Editorial: Go slow. Bulletin of the Atomic Scientists, 32(10):9, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Squ77].

Feld:1976:ENP

[Fel76e] Bernard T. Feld. Editorial: No peace for a hungry world. Bulletin of the Atomic Scientists, 32(7):8, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:ERT

[Fel76f] Bernard T. Feld. Editorial: Random thoughts for the new year. Bulletin of the Atomic Scientists, 32(1):4, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:ESS

[Fel76g] Bernard T. Feld. Editorial: Secrecy and security. Bulletin of the Atomic Scientists, 32(2):5–6, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:ETN

[Fel76h] Bernard T. Feld. Editorial: Technology and nuclear weapons. Bulletin of the Atomic Scientists, 32(5):6–7, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1976:SBD

[Fel76i] Bernard T. Feld. Slipping back into the Dark Ages. Bulletin of the Atomic Scientists, 32(6):3, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Whi76b].

Feld:1977:ETS

[Fel77a] Bernard T. Feld. Editorial: 1976–1977: time for a shock. Bulletin of the Atomic Scientists, 33(1):8–9, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SF77].

Feld:1977:EFS

[Fel77b] Bernard T. Feld. Editorial: A far-sighted energy program. Bulletin of the Atomic Scientists, 33(6):9, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:EEP

[Fel77c] Bernard T. Feld. Editorial: Energy and politics. Bulletin of the Atomic Scientists, 33(10):12–13, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:EGN

[Fel77d] Bernard T. Feld. Editorial: Good news and bad. Bulletin of the Atomic Scientists, 33(7):10–11, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [WF77].

Feld:1977:EHN

[Fel77e] Bernard T. Feld. Editorial: Hard-nosed by soft-headed. Bulletin of the Atomic Scientists, 33(4):9, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MM77].

Feld:1977:EJJ

[Fel77f] Bernard T. Feld. Editorial: Jaw, jaw, jaw. Bulletin of the Atomic Scientists, 33(10):13, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:ELP

[Fel77g] Bernard T. Feld. Editorial: Let’s not panic prematurely. Bulletin of the Atomic Scientists, 33(5):8–9, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho77].

Feld:1977:EOD

[Fel77h] Bernard T. Feld. Editorial: Ode to diversity. Bulletin of the Atomic Scientists, 33(8):10–11, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:ESS

[Fel77i] Bernard T. Feld. Editorial: SALT stalemate. Bulletin of the Atomic Scientists, 33(9):7, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:EWB

[Fel77j] Bernard T. Feld. Editorial: The way to begin is to stop. Bulletin of the Atomic Scientists, 33(3):9, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sol77].

Feld:1977:ENC

[Fel77k] Bernard T. Feld. Editor’s note: [civil defense]. Bulletin of the Atomic Scientists, 33(7):49, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:IC

[Fel77l] Bernard T. Feld. An international challenge. Bulletin of the Atomic Scientists, 33(3):5, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:NB

[Fel77m] Bernard T. Feld. The neutron bomb. Bulletin of the Atomic Scientists, 33(7):11, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1977:NDR

[Fel77n] Bernard T. Feld. The nuclear dilemma revisited. Bulletin of the Atomic Scientists, 33(2):7, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dud77b].

Feld:1977:BRY

[Fel77o] M. D. Feld. Books: Riga and Yalta: the roots of Soviet–American relations: The Shattered Peace: The Origins of the Cold War and the National Security State, by Daniel Yergin. Bulletin of the Atomic Scientists, 33(10):58–60, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:AME

[Fel78a] Bernard T. Feld. Arms and the Middle East. Bulletin of the Atomic Scientists, 34(2):38, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:DD

[Fel78b] Bernard T. Feld. The dilemma of deterrence. Bulletin of the Atomic Scientists, 34(3):12–13, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:EAS

[Fel78c] Bernard T. Feld. Editorial: Ambivalent steps toward a disarmed world. Bulletin of the Atomic Scientists, 34(4):8–9, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ESU

[Fel78d] Bernard T. Feld. Editorial: As others see us. Bulletin of the Atomic Scientists, 34(6):11, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ENPb

[Fel78e] Bernard T. Feld. Editorial: Norwegian prophecy or a gamble for peace? Bulletin of the Atomic Scientists, 34(10):5, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ERE

[Fel78f] Bernard T. Feld. Editorial: Recognizing the enemy. Bulletin of the Atomic Scientists, 34(7):4–5, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [EF78].

Feld:1978:ESH

[Fel78g] Bernard T. Feld. Editorial: SALT — a hostage to politics. Bulletin of the Atomic Scientists, 34(9):5, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ESS

[Fel78h] Bernard T. Feld. Editorial: Special Session in retrospect. Bulletin of the Atomic Scientists, 34(8):4, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:EMM

[Fel78i] Bernard T. Feld. Editorial: To move or not to move the clock. Bulletin of the Atomic Scientists, 34(1):8–9, January 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:EWP

[Fel78j] Bernard T. Feld. Editorial: Where is the present danger? Bulletin of the Atomic Scientists, 34(2):8, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ENH

[Fel78k] Bernard T. Feld. Editor’s note: [horizontal nuclear weapons proliferation]. Bulletin of the Atomic Scientists, 34(8):60, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ENPa

[Fel78l] Bernard T. Feld. Editor’s note: Pugwash beginnings. Bulletin of the Atomic Scientists, 34(4):35, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1978:ENS

[Fel78m] Bernard T. Feld. Editor’s note: SALT IIl. Bulletin of the Atomic Scientists, 34(6):32, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:CWS

[Fel79a] Bernard T. Feld. Can we survive a nuclear war with the Soviet Union? Bulletin of the Atomic Scientists, 35(7):38, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Goe80].

Feld:1979:DF

[Fel79b] Bernard T. Feld. Dear friends. Bulletin of the Atomic Scientists, 35(6):0, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EPD

[Fel79c] Bernard T. Feld. Editorial: A progressive decision. Bulletin of the Atomic Scientists, 35(9):4–5, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79i].

Feld:1979:EDT

[Fel79d] Bernard T. Feld. Editorial: Disaster through the back door. Bulletin of the Atomic Scientists, 35(5):4–5, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EEP

[Fel79e] Bernard T. Feld. Editorial: Environment in perspective. Bulletin of the Atomic Scientists, 35(8):5, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EGO

[Fel79f] Bernard T. Feld. Editorial: Great obsfuscations. Bulletin of the Atomic Scientists, 35(6):5, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:ENS

[Fel79g] Bernard T. Feld. Editorial: No substitute for SALT. Bulletin of the Atomic Scientists, 35(7):7, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EPT

[Fel79h] Bernard T. Feld. Editorial: Plain talk about nuclear proliferation. Bulletin of the Atomic Scientists, 35(1):4–5, January 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EBS

[Fel79i] Bernard T. Feld. Editorial: The Progressive secret. Bulletin of the Atomic Scientists, 35(5):5–6, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [San79, Han79, MSDP80] and outcome [Fel79c].

Feld:1979:EHP

[Fel79j] Bernard T. Feld. Editorial: The high price of SALT. Bulletin of the Atomic Scientists, 35(2):4, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EUT

[Fel79k] Bernard T. Feld. Editorial: The uneasy triangle. Bulletin of the Atomic Scientists, 35(4):4–5, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:ETM

[Fel79l] Bernard T. Feld. Editorial: Three Mile Island. Bulletin of the Atomic Scientists, 35(5):6, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Nus79, Fai79].

Feld:1979:EWH

[Fel79m] Bernard T. Feld. Editorial: What happened to the emperor? Bulletin of the Atomic Scientists, 35(9):4, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1979:EPN

[Fel79n] Bernard T. Feld. Einstein and the politics of nuclear weapons. Bulletin of the Atomic Scientists, 35(3):5–16, March 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Feld:1979:LKO

[Fel79o] Bernard T. Feld. Lew Kowarski, 1907–1979. Otto Robert Frisch, 1904–1979. Bulletin of the Atomic Scientists, 35(10):4–5, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:CCP

[Fel80a] Bernard T. Feld. Commentary: A clear and present choice. Bulletin of the Atomic Scientists, 36(7):5, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:EMM

[Fel80b] Bernard T. Feld. Editorial: 7 minutes to midnight: The hands move closer to midnight. Bulletin of the Atomic Scientists, 36(1): 1–3, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mat80a].

Feld:1980:ECW

[Fel80c] Bernard T. Feld. Editorial: Cold warriors take over. Bulletin of the Atomic Scientists, 36(3):4, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:EII

[Fel80d] Bernard T. Feld. Editorial: If it isn’t one thing it’s another! Bulletin of the Atomic Scientists, 36(7):4, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:ELA

[Fel80e] Bernard T. Feld. Editorial: Let’s agree — no first use! Bulletin of the Atomic Scientists, 36(5):6–7, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:EMT

[Fel80f] Bernard T. Feld. Editorial: Madder than MAD. Bulletin of the Atomic Scientists, 36(8):5, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:EMN

[Fel80g] Bernard T. Feld. Editorial: On the morning of November 5. Bulletin of the Atomic Scientists, 36(10):4, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:ESN

[Fel80h] Bernard T. Feld. Editorial: Setting up the next round. Bulletin of the Atomic Scientists, 36(2):4, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:ETS

[Fel80i] Bernard T. Feld. Editorial: Time to speak up. Bulletin of the Atomic Scientists, 36(6):5, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:ETD

[Fel80j] Bernard T. Feld. Editorial: Titan: dangerous and obsolete. Bulletin of the Atomic Scientists, 36(9):4, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:ENP

[Fel80k] Bernard T. Feld. Editors’ note: Pershing II: the Army’s strategic weapon. Bulletin of the Atomic Scientists, 36(8):25, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1980:SRN

[Fel80l] Bernard T. Feld. Special report: On non-proliferation policies and the NPT treaty. Bulletin of the Atomic Scientists, 36(4):37, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:BRB

[Fel81a] Bernard T. Feld. Book review: Reminiscences of Los Alamos, 1943–1945, edited by L. Badash, J. O. Hirschfelder, and H. P. Broida. Bulletin of the Atomic Scientists, 37(3):48, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:CIN

[Fel81b] Bernard T. Feld. [Comment on ‘Issues and non-issues in the nuclear policy debate’]. Bulletin of the Atomic Scientists, 37(10): 47, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:CP

[Fel81c] Bernard T. Feld. [Comment on Pugwash 1981]. Bulletin of the Atomic Scientists, 37(10):53, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:EMM

[Fel81d] Bernard T. Feld. Editorial: 4 minutes to midnight: The hands move closer to midnight. Bulletin of the Atomic Scientists, 37 (1):1, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SM81].

Feld:1981:ECLb

[Fel81e] Bernard T. Feld. Editorial: Campaign for a livable world. Bulletin of the Atomic Scientists, 37(6):1, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho81b].

Feld:1981:ECLa

[Fel81f] Bernard T. Feld. Editorial: Carter’s legacy. Bulletin of the Atomic Scientists, 37(3):5, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:ESS

[Fel81g] Bernard T. Feld. Editorial: Suicidal security. Bulletin of the Atomic Scientists, 37(2):4–5, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:EMEa

[Fel81h] Bernard T. Feld. Editorial: The means to the END. Bulletin of the Atomic Scientists, 37(5):9, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:EMEb

[Fel81i] Bernard T. Feld. Editorial: The Middle East fuse. Bulletin of the Atomic Scientists, 37(7):4, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:ENB

[Fel81j] Bernard T. Feld. Editorial: The neutron bomb decision. Bulletin of the Atomic Scientists, 37(8):5, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:ETT

[Fel81k] Bernard T. Feld. Editorial: Through a telescope backwards. Bulletin of the Atomic Scientists, 37(4):9, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:EWW

[Fel81l] Bernard T. Feld. Editorial: Words — not war. Bulletin of the Atomic Scientists, 37(10):4, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1981:GWS

[Fel81m] Bernard T. Feld. Gertrud Weiss Szilard, 1909–1981. Bulletin of the Atomic Scientists, 37(10):56, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:ATD

[Fel82a] Bernard T. Feld. After three decades. Bulletin of the Atomic Scientists, 38(4):2, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:EVJ

[Fel82b] Bernard T. Feld. Editorial: A visit to Japan. Bulletin of the Atomic Scientists, 38(2):2, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:EEH

[Fel82c] Bernard T. Feld. Editorial: Europe as ‘honest broker’. Bulletin of the Atomic Scientists, 38(3):2, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:ECS

[Fel82d] Bernard T. Feld. Editorial: The clock stands still. Bulletin of the Atomic Scientists, 38(1):3, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:MF

[Fel82e] Bernard T. Feld. A mutual freeze. Bulletin of the Atomic Scientists, 38(5):2–3, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:NNS

[Fel82f] Bernard T. Feld. New nuclear strategy? Bulletin of the Atomic Scientists, 38(4):3, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:NPP

[Fel82g] Bernard T. Feld. Nobel Peace Prize: [for 1982 to Alva R. Myrdal and Alfonso Garcia Robles]. Bulletin of the Atomic Scientists, 38(10):81, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:NTL

[Fel82h] Bernard T. Feld. [Note on “A Talk with Louis Harris”]. Bulletin of the Atomic Scientists, 38(7):3, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:PES

[Fel82i] Bernard T. Feld. Pugwash on European security. Bulletin of the Atomic Scientists, 38(7):43, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:WT

[Fel82j] Bernard T. Feld. Walking the tightrope. Bulletin of the Atomic Scientists, 38(9):2, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1982:WP

[Fel82k] Bernard T. Feld. A warning — not a prophecy. Bulletin of the Atomic Scientists, 38(6):2, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PF82].

Feld:1982:YA

[Fel82l] Bernard T. Feld. The year of appeals. Bulletin of the Atomic Scientists, 38(10):6–9, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mas83].

Feld:1983:BRB

[Fel83a] Bernard T. Feld. Book review: The Atomic Complex: a Worldwide Political History of Nuclear Energy, by Bertrand Goldschmidt. Bulletin of the Atomic Scientists, 39(1):37, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1983:EOP

[Fel83b] Bernard T. Feld. Editorial: Orwell’s prophecy. Bulletin of the Atomic Scientists, 39(3):2, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1983:EE

[Fel83c] Bernard T. Feld. End of an era. Bulletin of the Atomic Scientists, 39(1):2, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feliu:1983:BRB

[Fel83d] Alfred G. Feliu. Book reviews: “Do It My Way Or You’re Fired!”: Employee Rights and the Changing Role of Management Prerogatives, by David W. Ewing; Whistleblowing in Biomedical Research: Policies and Procedures for Responding to Reports of Misconduct. Bulletin of the Atomic Scientists, 39(10):40–42, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fels:1983:BRB

[Fel83e] Margaret F. Fels. Book review: The United States Energy Atlas, by David J. Cuff and William J. Young. Bulletin of the Atomic Scientists, 39(2):40–41, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:AP

[Fel84a] Bernard T. Feld. Agenda 1984: Perspectives for 1984. Bulletin of the Atomic Scientists, 40(1):3, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:BNB

[Fel84b] Bernard T. Feld. Book note: Energy and Order: Some Reflections on Evolution, by Charles Price. Bulletin of the Atomic Scientists, 40(3):47, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:EES

[Fel84c] Bernard T. Feld. Editorial: End space race now. Bulletin of the Atomic Scientists, 40(8):2, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:ESI

[Fel84d] Bernard T. Feld. Editorial: Step by inexorable step. Bulletin of the Atomic Scientists, 40(2):2, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho84b].

Feld:1984:ETM

[Fel84e] Bernard T. Feld. Editorial: Three minutes to midnight. Bulletin of the Atomic Scientists, 40(1):2, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mas84].

Feld:1984:PK

[Fel84f] Bernard T. Feld. Pyotr L. Kapitsa, 1894–1984. Bulletin of the Atomic Scientists, 40(6):2, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:TRA

[Fel84g] Bernard T. Feld. Tribute to Ruth Adams. Bulletin of the Atomic Scientists, 40(9):2, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1984:V

[Fel84h] Bernard T. Feld. Valedictory. Bulletin of the Atomic Scientists, 40(10):3–4, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1985:FYM

[Fel85a] Bernard T. Feld. Forty years of muddling through. Bulletin of the Atomic Scientists, 41(7):30–33, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1985:LLH

[Fel85b] Bernard T. Feld. Letter: Lessons from Hiroshima and Nagasaki. Bulletin of the Atomic Scientists, 41(11):65–66, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1985:SWV

[Fel85c] Bernard T. Feld. Star Wars not viable. Bulletin of the Atomic Scientists, 41(4):60–61, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1986:BRB

[Fel86a] Bernard T. Feld. Book review: Bird of Passage, by Rudolf Peierls. Bulletin of the Atomic Scientists, 42(5):44, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fellman:1986:LAF

[Fel86b] Gordon Fellman. Letter: African famine. Bulletin of the Atomic Scientists, 42(1):54–55, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1988:LBC

[Fel88] Bernard T. Feld. Letter: Ban the cruise. Bulletin of the Atomic Scientists, 44(7):60–61, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mac88].

Feld:1992:BRL

[Fel92] Bernard T. Feld. Book review: A life in science: A Different Sort of Time: The Life of Jerrold Zacharias, Scientist, Engineer, Educator, by Jack S. Golstein. Bulletin of the Atomic Scientists, 48(8):43–44, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feld:1993:CWT

[Fel93] Bernard T. Feld. “Cold warriors take over,” a March 1980 editorial. Bulletin of the Atomic Scientists, 49(4):14, May 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feldhoff:2014:PFE

[Fel14] Thomas Feldhoff. Post-Fukushima energy paths: Japan and Germany compared. Bulletin of the Atomic Scientists, 70(6):87–96, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fenstermacher:1991:ARN

[Fen91] Dan Fenstermacher. Arms race: The next generation. Bulletin of the Atomic Scientists, 47(2):29–33, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ferretti:1952:SOE

[Fer52] Bruno Ferretti. Some other European experiences. Bulletin of the Atomic Scientists, 8(7):247–248, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Fermi:1953:BBP

[Fer53] Laura Fermi. Books: Moral Principles of Action, planned and edited by Ruth Nanda Anshen. Bulletin of the Atomic Scientists, 9(8):316–317, 320, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1954:LAL

[Fer54a] Enrico Fermi. Letter: Los Alamos Laboratory has deserved the gratitude of this nation. Bulletin of the Atomic Scientists, 10(9): 359, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1954:BBC

[Fer54b] Laura Fermi. Books: The Captive Mind, by Czeslaw Milosz. Bulletin of the Atomic Scientists, 10(4):139, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1957:BBM

[Fer57a] Laura Fermi. Books: Science and Modern Life, by Sir E. John Russell; Science in Our Lives, by Ritchie Calder; Facing the Atomic Future, by E. W. Titterton. Bulletin of the Atomic Scientists, 13(2):72, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1957:BBN

[Fer57b] Laura Fermi. Books: The New World of the Atom, by James Stokley. Bulletin of the Atomic Scientists, 13(3):109, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1957:LEN

[Fer57c] Laura Fermi. Letter to the Editor: The nature of nuclear warfare. Bulletin of the Atomic Scientists, 13(6):232, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel57b].

Ferry:1960:LEA

[Fer60] W. H. Ferry. Letter to the Editor: Alternatives to arms race. Bulletin of the Atomic Scientists, 16(4):111, 145, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ferry:1961:SRC

[Fer61] W. H. Ferry. Some replies on China. Bulletin of the Atomic Scientists, 17(8):346, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1962:BBI

[Fer62] Lauri Fermi. Books: The Irreversible Decision, 1939–1950, by Robert C. Batchelder. Bulletin of the Atomic Scientists, 18(5): 38–39, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ferry:1963:WPP

[Fer63] W. H. Ferry. What price peace? Bulletin of the Atomic Scientists, 19(7):19–23, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1968:AFF

[Fer68] Laura Fermi. After the fall of France: The Emergency Rescue Committee. Bulletin of the Atomic Scientists, 24(2):9–13, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1969:BBC

[Fer69a] Laura Fermi. Books: The Civilizing Mission: A History of the Italo–Ethiopian War, by A. J. Barker. Bulletin of the Atomic Scientists, 25(5):51–52, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1969:CAP

[Fer69b] Laura Fermi. Cars and air pollution. Bulletin of the Atomic Scientists, 25(8):35–37, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1970:BR

[Fer70] Laura Fermi. Bombs or reactors? Bulletin of the Atomic Scientists, 26(6):28–29, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ferry:1987:LPS

[Fer87] Phyllis C. Ferry. Letter: Pentagon subsidizes Bulletin? Bulletin of the Atomic Scientists, 43(9):54, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ferguson:1988:DAS

[Fer88] Bruce Ferguson. Different agendas, styles shape SANE/freeze. Bulletin of the Atomic Scientists, 44(3):26–30, April 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [LC88].

Ferguson:1999:TTG

[Fer99] Charles D. Ferguson. Tritium: TVA gets the nod. Bulletin of the Atomic Scientists, 55(2):12–14, March 1999. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/2/12.

Ferguson:2003:PM

[Fer03] Charles D. Ferguson. Policy by E-Mail. Bulletin of the Atomic Scientists, 59(2):70–72, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/59/2/70.

Ferguson:2010:NCP

[Fer10] Charles D. Ferguson. Next customer, please: The risk in conventional arms sales along with nuclear energy deals. Bulletin of the Atomic Scientists, 66(6):36–42, November/December 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/6/36.

Fessenden:1973:CSH

[Fes73] Frederic P. Fessenden. Communications: Shoreham hearings. Bulletin of the Atomic Scientists, 29(1):2, January 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano72g].

Festag:1976:LNF

[Fes76] Eckhard Festag. Letter: Nuclear freeze. Bulletin of the Atomic Scientists, 32(2):2, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sal75b].

Festag:1977:LCU

[Fes77] Eckhard Festag. Letter: Conserving uranium. Bulletin of the Atomic Scientists, 33(10):4–5, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet77a, Maw78b].

Fetter:1987:WTB

[Fet87] Steve Fetter. Would a test ban strengthen SDI? Bulletin of the Atomic Scientists, 43(9):40–42, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fetter:2000:E

[Fet00] Steve Fetter. Energy 2050. Bulletin of the Atomic Scientists, 56 (4):28–38, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/4/28.

Ford:2014:EW

[FG14] Kenneth Ford and Clark Glymour. The enhanced warfighter. Bulletin of the Atomic Scientists, 70(1):43–53, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/70/1/43.
full.pdf+html.

Fallers:1965:PPN

[FGJ65] Lloyd A. Fallers, Clifford Geertz, and Morris Janowitz. The policy proposals: “A negotiated stalemate”. Bulletin of the Atomic Scientists, 21(6):42–45, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fainberg:1983:CCD

[FH83] Anthony Fainberg and John P. Holdren. Commentary: ‘The connection is dangerous’. Bulletin of the Atomic Scientists, 39(5): 60–62, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol83b].

Franck:1946:RSW

[FHN^{+}46] James Franck, Donald J. Hughes, J. J. Nickson, Eugene Rabinowitch, Glenn T. Seaborg, J. C. Stearns, and Leo Szilard. A report to the Secretary of War. Bulletin of the Atomic Scientists, 1(10):2–4, 16, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:/

/datatorch.com/Science/Scientists_Stories.aspx?id=52;
 http://en.wikipedia.org/wiki/Franck_Report;   http:/
/nuclearsecrecy.com/blog/2012/01/11/weekly-document-9-
the-uncensored-franck-report-1945-1946/;   http://www.
atomicarchive.com/Docs/ManhattanProject/FranckReport.
shtml;   http://www.dannen.com/decision/franck.html;
http://www.fas.org/sgp/eprint/franck.html;   http://www.
nuclearfiles.org/menu/key-issues/ethics/issues/scientific/
franck-report.htm;   http://www.wissenschaft-und-frieden.
de/seite.php?artikelID=1092. The authors are members of

the Committee on Political and Social Problems, Manhattan Project ‘Metallurgical Laboratory’, University of Chicago, and this report subsequently became known as ‘The Franck Report’. The report was written on 11 June 1945, one month before the first successful test of the atomic bomb at the Trinity site in New Mexico. In it, seven scientists, most of who worked on the Manhattan Project, recommend against the use of the atomic bomb on Japan, and predict a dangerous post-war nuclear arms race. This version was partially declassified in February 1946; eight fragments of text have been removed at the order of government censors. Some of that censored text was later recovered, but at least one sentence may still be missing — see the Web document “The Uncensored Franck Report (1945–1946)”, by Alex Wellerstein, dated 11 January 2012. Reprinted in [FHN^{+}63], and in a longer form, in [Smi65, Appendix] and [Smi70b, pages 371–383].

Franck:1963:RSW

[FHN^{+}63] James Franck, Donald J. Hughes, I. I. Nickson, Eugene Rabinowitch, Glenn T. Seaborg, Joyce C. Stearns, and Leo Szilard. A report to the Secretary of War. In Grodzins and Rabinowitch [GR63], pages 19–27. LCCN D842 .B78. Reprint of [FHN^{+}46].

Field:1978:LPA

[Fie78] Sylvia S. Field. Letter: Pointing to alternatives. Bulletin of the Atomic Scientists, 34(5):6, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Par78b].

Fieldhouse:1987:ARS

[Fie87] Richard Fieldhouse. Arms race at sea: Nuclear weapons at sea. Bulletin of the Atomic Scientists, 43(7):19–23, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fieldhouse:1990:CMC

[Fie90] Richard Fieldhouse. Cruise missile compromise surfacing. Bulletin of the Atomic Scientists, 46(5):21–24, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fieldhouse:1991:CMS

[Fie91] Richard Fieldhouse. China’s mixed signals on nuclear weapons. Bulletin of the Atomic Scientists, 47(4):37–42, May 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Field:2018:SAE

[Fie18] Christopher B. Field. Smart adaptation in an era of rising climate risks. Bulletin of the Atomic Scientists, 74(2):73–80, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1436806.

Field:2019:WCS

[Fie19] Matt Field. The world can support far more trees. planting them can reduce carbon pollution a lot: An interview with professor Tom Crowther. Bulletin of the Atomic Scientists, 75 (5):236–238, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1654267.

Field:2021:HCB

[Fie21a] Matt Field. How can the Biden administration reduce scientific disinformation? Slow the high-pressure pace of scientific publishing. Bulletin of the Atomic Scientists, 77(1):38–40, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1860332.

Field:2021:IRF

[Fie21b] Matt Field. Instead of reforming Facebook, should we just build something else? Bulletin of the Atomic Scientists, 77 (3):116–118, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1912052.

Field:2024:RJP

[Fie24] Matt Field. RFK Jr.’s presidential ambitions may have fallen short, but his anti-vax beliefs are winning in many statehouses. Bulletin of the Atomic Scientists, 80(5):302–307, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2388466.

Figueres:2023:WMS

[Fig23] Christiana Figueres. Why a mind-set of stubborn optimism about the climate crisis is needed, now more than ever. Bulletin of the Atomic Scientists, 80(1):38–40, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2293576.

Filner:1976:RPA

[Fil76] Robert E. Filner. Roots of political activism in British science. Bulletin of the Atomic Scientists, 32(1):25–29, January 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finletter:1953:PO

[Fin53] Thomas K. Finletter. Priority one. Bulletin of the Atomic Scientists, 9(10):365, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finney:1954:AD

[Fin54a] Nat S. Finney. Atomic dilemma. Bulletin of the Atomic Scientists, 10(4):116–118, 124, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finney:1954:TAS

[Fin54b] Nat S. Finney. The threat to atomic science. Bulletin of the Atomic Scientists, 10(7):285–286, 295, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finney:1955:RVA

[Fin55] Nat S. Finney. A reporter’s views on atomic secrecy. Bulletin of the Atomic Scientists, 11(7):247–249, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finer:1956:PVS

[Fin56] Herman Finer. Part V: Science and government: Government and the expert. Bulletin of the Atomic Scientists, 12(9):329– 332, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finletter:1957:FD

[Fin57a] Thomas K. Finletter. Facing disarmament. Bulletin of the Atomic Scientists, 13(5):154–156, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finney:1957:BBP

[Fin57b] Nat S. Finney. Books: Atoms and People, by Ralph Lapp. Bulletin of the Atomic Scientists, 13(2):71, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finney:1957:LES

[Fin57c] Nat S. Finney. Letter to the Editor: The small band of atomic scientists. Bulletin of the Atomic Scientists, 13(6):231, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finkelstein:1965:BBK

[Fin65] Lawrence S. Finkelstein. Books: Peace-Keeping by the U.N. Forces: From Suez to the Congo, by Arthur Lee Burns and Nina Heathcote; International Military Forces: The Question of Peace-Keeping in an Armed and Disarming World. Bulletin of the Atomic Scientists, 21(5):36–37, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finlayson:1975:VO

[Fin75] Fred C. Finlayson. A view from the outside. Bulletin of the Atomic Scientists, 31(7):20–25, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fine:1977:RPR

[Fin77a] Jacob Fine. Rx: a peer review systems for physicians. Bulletin of the Atomic Scientists, 33(7):38–43, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gne77].

Fine:1977:SSC

[Fin77b] Jacob Fine. [Sidebar:] some common objections to medical peer review. Bulletin of the Atomic Scientists, 33(7):42, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fingar:1984:STCb

[Fin84] Thomas Fingar. Science and technology in China: Consequences of catching up. Bulletin of the Atomic Scientists, 40(8):13s–15s [various between 24–25], October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finston:1990:BRR

[Fin90] Roland Finston. Book review: Radiation without representation: Living with Radiation: The Risk, the Promise, by Henry N. Wagner, Jr. and Linda E. Ketchum; Multiple Exposures: Chronicles of the Radiation Age, by Catherine Caulfield. Bulletin of the Atomic Scientists, 46(7):41–42, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finston:1992:LPG

[Fin92] Roland A. Finston. Letter: Plutonium guinea pig likely to survive. Bulletin of the Atomic Scientists, 48(10):51, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot92q].

Findlay:2001:UP

[Fin01] Trevor Findlay. An unwitting pawn? Bulletin of the Atomic Scientists, 57(4):67–68, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/4/67.

Fisher:1946:AEC

[Fis46] Jerome Fisher. Atomic energy controls: 4. Aerial survey — an aid in mining control. Bulletin of the Atomic Scientists of Chicago, 1(5):1–2, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1955:BRB

[Fis55] Joachim Fischer. Book review: V-2, by Walter Dornberger. Bulletin of the Atomic Scientists, 11(3):95, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1961:FPA

[Fis61a] Louis Fischer. A foreign policy for America. Bulletin of the Atomic Scientists, 17(4):153–156, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fisher:1961:DR

[Fis61b] Roger Fisher. Deputy for restraint. Bulletin of the Atomic Scientists, 17(7):272–273, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fisher:1962:DWW

[Fis62a] Roger Fisher. Do we want to “Win” the Cold War? Bulletin of the Atomic Scientists, 18(1):33–35, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fisher:1962:RDV

[Fis62b] Roger Fisher. Responding to disarmament violations. Bulletin of the Atomic Scientists, 18(7):22–28, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fishman:1962:EPD

[Fis62c] Leslie Fishman. An economic plan for disarmament. Bulletin of the Atomic Scientists, 18(3):37–38, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1976:LUS

[Fis76] C. David Fischer, Jr. Letter: Uranium shortage. Bulletin of the Atomic Scientists, 32(7):4–5, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ner76].

Fisher:1981:CSF

[Fis81] Roger Fisher. Commentary from San Francisco: Preventing nuclear war. Bulletin of the Atomic Scientists, 37(3):11–17, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wal81b].

Fischer:1982:MIS

[Fis82] D. A. V. Fischer. More on IAEA safeguards. Bulletin of the Atomic Scientists, 38(5):62, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mog81].

Fischer:1983:BRB

[Fis83] Mary Ellen Fischer. Book reviews: The Domestic Context of Soviet Foreign Policy, edited by Seweryn Bialer; The Soviet Union in the Third World: Successes and Failures, edited by Robert H. Donaldson; Soviet Influence in Easter Europe: Political Autonomy and the Warsaw Pact, by Christopher D. Jones. Bulletin of the Atomic Scientists, 39(1):31–33, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1984:EEU

[Fis84] Mary Ellen Fischer. Easter Europe: the unstable buffer. Bulletin of the Atomic Scientists, 40(10):16–20, 21–22, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1986:CNS

[Fis86a] David A. V. Fischer. The challenge of nuclear safeguards. Bulletin of the Atomic Scientists, 42(6):29–33, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1986:CES

[Fis86b] David A. V. Fischer. Chernobyl: the emerging story: The international response. Bulletin of the Atomic Scientists, 42(7): 46–48, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1989:LTI

[Fis89] David Fischer. Letter: Timely inspections. Bulletin of the Atomic Scientists, 45(7):44–45, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hir89b] and comment [Hir89a].

Fischer:1990:NRE

[Fis90] David A. V. Fischer. NPT review: Eastern Europe after Pax Sovietica. Bulletin of the Atomic Scientists, 46(6):23, 26, July/ August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fisher:1991:PBC

[Fis91] Cathleen S. Fisher. Perspective: Build confidence, not weapons. Bulletin of the Atomic Scientists, 47(5):11, 42, June 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fitzgerald:1951:WNR

[Fit51] Dennis A. Fitzgerald. World needs and resources. Bulletin of the Atomic Scientists, 7(4):102–105, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fitzgerald:1966:DFP

[Fit66] C. P. Fitzgerald. The directions of foreign policy. Bulletin of the Atomic Scientists, 22(6):65–70, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fitzgerald:1969:RH

[Fit69] C. P. Fitzgerald. A revolutionary hiatus. Bulletin of the Atomic Scientists, 25(2):53–60, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Fitch:1975:AOT

[Fit75] Val L. Fitch. ‘All in our time”: View from bottom. Bulletin of the Atomic Scientists, 31(2):43–46, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fitzpatrick:2011:HEV

[Fit11] Mark Fitzpatrick. How Europeans view tactical nuclear weapons on their continent. Bulletin of the Atomic Scientists, 67(2): 57–65, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/2/57.

Farinella:1991:PJV

[FJ91] Paolo Farinella and Venance Journe. Perspective: Justice for Vanunu. Bulletin of the Atomic Scientists, 47(1):14, January/ February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischbach:2012:RDT

[FJ12] Ephraim Fischbach and Jere Jenkins. Radiation detection: There’s an app for that. Bulletin of the Atomic Scientists, 68(2): 63–69, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/2/63.

Ford:1971:TAS

[FK71] Norman C. Ford and Joseph W. Kane. 2. Technical alternatives: Solar power. Bulletin of the Atomic Scientists, 27(8):27–31, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1989:LII

[FK89] Jerome D. Frank and Tatiana Kabachenko. Letters: International incident. Bulletin of the Atomic Scientists, 45(2):47–48, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Funabashi:2012:FRC

[FK12] Yoichi Funabashi and Kay Kitazawa. Fukushima in review: a complex disaster, a disastrous response. Bulletin of the Atomic Scientists, 68(2):9–21, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/2/9.

Ferrari:1987:MAS

[FKM87] Paul L. Ferrari, Jeffrey W. Knopf, and Raul L. Madrid. Major arms suppliers to 12 Third World conflicts. Bulletin of the Atomic Scientists, 43(10):34, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fischer:1962:LEP

[FL62] Roland Fischer and E. James Lieberman. Letters to the Editor: Psychochemicals. Bulletin of the Atomic Scientists, 18(7): 36, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lie62b, LLH^{+}62].

Flavin:1996:NPB

[FL96] Christopher Flavin and Nicholas Lenssen. Nuclear power browning out. Bulletin of the Atomic Scientists, 52(3):52–55, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flamm:1960:LEN

[Fla60] Irving H. Flamm. Letter to the Editor: No waste necessary. Bulletin of the Atomic Scientists, 16(6):257, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pie60].

Flacks:1965:SSI

[Fla65] Richard Flacks. Some social implications of the teach-ins. Bulletin of the Atomic Scientists, 21(8):21, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flatte:1968:BBR

[Fla68] Stanley M. Flatté. Books: Symmetries and Reflections, by Eugene P. Wigner; The Relevance of Physics, by Stanley L. Jaki. Bulletin of the Atomic Scientists, 24(3):28–30, March 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flanagan:1978:CWH

[Fla78] Stephen J. Flanagan. Congress, the White House and SALT. Bulletin of the Atomic Scientists, 34(9):34–40, November 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flamberg:1980:QNI

[Fla80a] Daniel Flamberg. Query: [national intelligence estimates]. Bulletin of the Atomic Scientists, 36(3):64, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flannery:1980:FRB

[Fla80b] William Flannery. Film review: Oil and American Power. Bulletin of the Atomic Scientists, 36(5):56–57, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flavin:1987:NPB

[Fla87] Christopher Flavin. Nuclear power’s burdened future. Bulletin of the Atomic Scientists, 43(6):26–31, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gri87b].

Flanagan:1988:NDO

[Fla88] Stephen J. Flanagan. Nonoffensive defense is overrated. Bulletin of the Atomic Scientists, 44(7):46–48, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flanagan:1990:BRC

[Fla90] Dennis Flanagan. Book review: Cosmic odd couple From Quarks to the Cosmos: Tools of Discovery, by Leon M. Lederman and David N. Schramm. Bulletin of the Atomic Scientists, 46(2): 41–42, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fletcher:1958:LER

[Fle58] Allan L. Fletcher. Letter to the Editor: To relax the tension. Bulletin of the Atomic Scientists, 14(5):200–201, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fleischman:1959:LEB

[Fle59] R. Fleischman. Letter to the Editor: Berlin. Bulletin of the Atomic Scientists, 15(7):321, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab59a, Rab59c].

Fletcher:1970:PIG

[Fle70] Joseph O. Fletcher. Polar ice and the global climate machine. Bulletin of the Atomic Scientists, 26(10):40–47, December 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Forsberg:1991:CFT

[FLLW91a] Randall Forsberg, Rob Leavitt, and Steve Lilly-Weber. Conventional forces treaty buries Cold War. Bulletin of the Atomic Scientists, 47(1):32–37, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forsberg:1991:MT

[FLLW91b] Randall Forsberg, Rob Leavitt, and Steve Lilly-Weber. Musical tanks. Bulletin of the Atomic Scientists, 47(1):36, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forsberg:1991:UC

[FLLW91c] Randall Forsberg, Rob Leavitt, and Steve Lilly-Weber. Understanding CFE. Bulletin of the Atomic Scientists, 47(1):34, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flood:1976:NS

[Flo76a] Michael Flood. Nuclear sabotage. Bulletin of the Atomic Scientists, 32(8):29–36, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pen77b, Sut77].

Flowers:1976:WVN

[Flo76b] Sir Brian Flowers. A watchdog’s view: Nuclear power and the public interest. Bulletin of the Atomic Scientists, 32(10):24–27, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tor77b].

Flowers:1978:NPP

[Flo78] Brian Flowers. Nuclear power: a perspective of the risks, benefits and options. Bulletin of the Atomic Scientists, 34(3):21–26, 54– 57, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ung78b].

Flowers:1981:CGM

[Flo81] R. Wills Flowers. Commentary: On growth mania. Bulletin of the Atomic Scientists, 37(7):63, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap81a].

Flowers:1985:LUG

[Flo85] R. Wills Flowers. Letter: Untitled [global future]. Bulletin of the Atomic Scientists, 41(2):51, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim84d].

Flournoy:1988:WRI

[Flo88] Michele Flournoy. Washington report: INF: the first 60 days; air force may have twins; South Africa still in IAEA; Teller’s telltale letters; what the GAO didn’t let you read. Bulletin of the Atomic Scientists, 44(9):3–5, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flournoy:1989:WRN

[Flo89a] Michèle A. Flournoy. Washington report: Navy shuns credit for cuts; Bush chases review blues; Britain to get airborne INF. Bulletin of the Atomic Scientists, 45(6):3–4, July/August 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flournoy:1989:WRR

[Flo89b] Michèle A. Flournoy. Washington report: Ripe for revision; changing ACDA’s act; brass-to-brass; what GLCMs? Bulletin of the Atomic Scientists, 45(4):4–5, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Florini:1990:RVC

[Flo90] Ann M. Florini. Report: Verification: Come spy the friendly skies. Bulletin of the Atomic Scientists, 46(2):13, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1997:BPU

[Fly97] Michael Flynn. Bulletin: Prospects of U.S. sales fuels Latin “jet gap” fears. Bulletin of the Atomic Scientists, 53(6):5–7, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1998:BRR

[Fly98a] Michael Flynn. Book review: Not a religious war: God and the Gun: The Church and Irish Terrorism, by Martin Dillon. Bulletin of the Atomic Scientists, 54(4):57–59, July/August 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1998:BPF

[Fly98b] Michael Flynn. Bulletin: “A perfect failure”. Bulletin of the Atomic Scientists, 54(3):7–9, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1998:BDT

[Fly98c] Michael Flynn. Bulletin: Despite talk, Juraguá only a maybe. Bulletin of the Atomic Scientists, 54(1):6–7, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1999:DIY

[Fly99a] Michael Flynn. The do-it-yourself reactor. Bulletin of the Atomic Scientists, 55(1):7, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1999:PM

[Fly99b] Michael Flynn. Political minefield. Bulletin of the Atomic Scientists, 55(2):49–53, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/2/49.

Flynn:1999:WDY

[Fly99c] Michael Flynn. What did you do in the cold war, daddy? Bulletin of the Atomic Scientists, 55(5):9, September/October 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:2000:BSB

[Fly00a] Michael Flynn. Bernardo Salas, “Bad Element”. Bulletin of the Atomic Scientists, 56(4):45–51, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/4/45.

Flynn:2000:CNP

[Fly00b] Michael Flynn. Chiapas: No place like home. Bulletin of the Atomic Scientists, 56(5):46–51, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/46.

Flynn:2000:MTG

[Fly00c] Michael Flynn. Mexico: Trouble at The Green Lagoon. Bulletin of the Atomic Scientists, 56(3):12–14, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/3/12.

Flynn:2000:PF

[Fly00d] Michael Flynn. Playing with fire. Bulletin of the Atomic Scientists, 56(5):35–36, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/5/35.

Flynn:2000:WSC

[Fly00e] Michael Flynn. “We served the cause”. Bulletin of the Atomic Scientists, 56(5):41–44, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/41.

Flynn:2001:BNR

[Fly01a] Matthew Flynn. Brazil: Nuclear to the rescue? Bulletin of the Atomic Scientists, 57(5):15–17, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/15.

Flynn:2001:DLO

[Fly01b] Michael Flynn. A debt Long overdue. Bulletin of the Atomic Scientists, 57(4):38–48, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/4/38.

Flynn:2001:RGU

[Fly01c] Michael Flynn. Russia: Guilty until proven innocent. Bulletin of the Atomic Scientists, 57(2):14–17, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/2/14.

Flynn:2002:DEF

[Fly02a] Michael Flynn. ‘?Dónde Está La Frontera? Bulletin of the Atomic Scientists, 58(4):24–35, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/24.

Flynn:2002:P

[Fly02b] Michael Flynn. Is this peace? Bulletin of the Atomic Scientists, 58(6):62–70, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/62.

Flynn:2002:RSM

[Fly02c] Michael Flynn. Russia: Spy mania. Bulletin of the Atomic Scientists, 58(5):15–16, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/5/15.

Flynn:2002:SSH

[Fly02d] Michael Flynn. Searching for safe haven. Bulletin of the Atomic Scientists, 58(6):22–25, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/6/22.

Flynn:2003:BAS

[Fly03] Matthew Flynn. Brazil: The Amazon’s silent spy. Bulletin of the Atomic Scientists, 59(1):12–14, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/1/12.

Flynn:2005:EWD

[Fly05a] Michael Flynn. Ecuador: What’s the deal at Manta? Bulletin of the Atomic Scientists, 61(1):23–29, January/February 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/1/23.

Flynn:2005:LCW

[Fly05b] Michael Flynn. Landmines: Clearing the way. Bulletin of the Atomic Scientists, 61(5):18–19, September/October 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/5/18.

Flynn:2006:CHL

[Fly06a] Michael Flynn. Chernobyl: Hardly the last word. Bulletin of the Atomic Scientists, 62(2):14–16, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/2/14.

Flynn:2006:EBN

[Fly06b] Michael Flynn. Europe: On its borders, new problems. Bulletin of the Atomic Scientists, 62(6):21–23, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/6/21.

Fischer:1982:CIS

[FM82] David A. V. Fischer and Sidney Moglewer. Commentary: On IAEA safeguards. Bulletin of the Atomic Scientists, 38(3):39–42, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mog81].

Forge:1993:LMN

[FM93] John Forge and Sverre Myhra. Letter: Mythic nukes? Bulletin of the Atomic Scientists, 49(6):54–55, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alb93e].

Findlay:2001:NTV

[FM01] Trevor Findlay and Oliver Meier. Nuclear testing: In verification we trust. Bulletin of the Atomic Scientists, 57(1):13–15, January/ February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
1/13.

Ferguson:2007:MIB

[FO07] Charles D. Ferguson and Lisa Obrentz. Make it or break it. Bulletin of the Atomic Scientists, 63(2):??, March/April 2007. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/63/2/46.

Foerstner:2007:WVA

[Foe07] Abigail Foerstner. What Van Allen found in space. Bulletin of the Atomic Scientists, 63(4):??, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/58.

Foley:1960:LEB

[Fol60] G. M. Foley. Letter to the Editor: The bridge at Geneva. Bulletin of the Atomic Scientists, 16(8):345, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Foley:1974:BBC

[Fol74] Ann D. Foley. Books: The Curve of Binding Energy: A Journey into the Awesome and Alarming World of Theodore B. Taylor, by John McPhee. Bulletin of the Atomic Scientists, 30(8):49, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Foley:1975:BBV

[Fol75] Ann D. Foley. Books: The Voyages of Apollo: The Exploration of the Moon, by Richard S. Lewis. Bulletin of the Atomic Scientists, 31(4):48, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forrestal:1949:SFS

[For49] James Forrestal. Secretary Forrestal’s statement on biological warfare. Bulletin of the Atomic Scientists, 5(4):104–105, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forrestal:1951:EBFa

[For51a] James Vincent Forrestal. Excerpts from The Forrestal Diaries: “The Story of a Nation’s Deepening Crisis”. [Part I]. Bulletin of the Atomic Scientists, 7(10):292–296, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forrestal:1951:EBFb

[For51b] James Vincent Forrestal. Excerpts from The Forrestal Diaries: “The Story of a Nation’s Deepening Crisis”: Part II. Bulletin of the Atomic Scientists, 7(11):327–331, 344, November 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forbes:1957:RSA

[For57] William H. Forbes. The role of science attachés. Bulletin of the Atomic Scientists, 13(8):274–276, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forscher:1964:LES

[For64] Frederick Forscher. Letter to the Editor: Science. Bulletin of the Atomic Scientists, 20(10):23, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [She64c].

Forland:1970:CLA

[For70] Katrine Seip Forland. Communications: Letter from Africa: Higher education in East Africa. Bulletin of the Atomic Scientists, 26(4):42–45, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forscher:1975:LTT

[For75] Frederick Forscher. Letters: A third truth. Bulletin of the Atomic Scientists, 31(6):2–3, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste75b].

Forman:1980:BRE

[For80] Paul Forman. Book reviews: Einstein: A Centenary Volume, edited by A. P. French; Albert Einstein’s Theory of General Relativity, edited by Gerald E. Tauber; Einstein, by Louis de Broglie and others; Einstein. Sein Leben und seine Zeit, by Philipp Frank; Documentary publications: Albert Einstein Autobiographical Notes, translated and edited by Paul Arthur Schilpp; Elie Cartan, Albert Einstein, Letters on Absolute Parallelism 1929– 1932, edited by Robert Debever; Aus meinen späten Jahren, by Albert Einstein; Albert Einstein in Berlin 1918–1933, Teil I: Darstellung und Dokumente; Teil II: Spezialinventar; Images of Einstein: a Catalog, compiled by Joan N. Warnow; Albert Einstein: the Human Side. New Glimpses from his Archives, edited by Helen Dukas and Banesh Hoffman; Einstein privat. Herta W. erinnert sich an die Jahre 1927 bis 1933, by Friedrich Herneck; Centenary exhibitions: Albert Einstein, 1879–1955. A Centenary Exhibit of Manuscripts, Books, and Portraits Selected from the Humanities Research Center Collections, by Albert C. Lewis; Gedächtnisausstellung zum 100. Geburtstag von Albert Einstein, Otto Hahn, Max von Laue, Lise Meitner; Einstein = mc^{2}; Einstein 1879–1979. Exhibition. Bulletin of the Atomic Scientists, 36(3):49–55, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ford:1983:BRB

[For83] Daniel F. Ford. Book review: Living with Nuclear Weapons, by the Harvard Nuclear Study Group. Bulletin of the Atomic Scientists, 39(8):34, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forsberg:1985:PCN

[For85] Randall Forsberg. Parallel cuts in nuclear and conventional forces. Bulletin of the Atomic Scientists, 41(7):152–156, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forman:1986:BRB

[For86] Paul Forman. Book review: By the Bomb’s Early Light: American Thought and Culture at the Dawn of the Atomic Age, by Paul Boyer. Bulletin of the Atomic Scientists, 42(4):48–49, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Boy86b, Far86].

Ford:1987:BRB

[For87] Daniel Ford. Book review: Managing Nuclear Operations, edited by Ashton B. Carter, John D. Steinbruner, and Charles A. Zraket. Bulletin of the Atomic Scientists, 43(10):43–45, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forsberg:1988:TNW

[For88] Randall Forsberg. Toward a nonaggressive world. Bulletin of the Atomic Scientists, 44(7):49–54, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forsberg:1991:SEA

[For91] Randall Forsberg. START: End arms control, begin disarmament. Bulletin of the Atomic Scientists, 47(9):29–31, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forsberg:1992:KPP

[For92] Randall Forsberg. Keep peace by pooling armies. Bulletin of the Atomic Scientists, 48(4):41–42, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forden:2002:LDW

[For02] Geoffrey Forden. Laser defenses: What if they work? Bulletin of the Atomic Scientists, 58(5):48–53, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/5/48.

Forsyth:2006:DM

[For06] James Forsyth. Dangerous minds? Bulletin of the Atomic Scientists, 62(3):56–57, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/3/56.

Forsberg:2009:RPG

[For09] Charles Forsberg. The real path to Green energy: Hybrid nuclearrenewable power. Bulletin of the Atomic Scientists, 65(6):65–71, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/65/
6/65.

Forsberg:2020:KPP

[For20] Randall Forsberg. 1992: Keep peace by pooling armies. Bulletin of the Atomic Scientists, 76(6):395–396, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847499.

Foster:1988:CJN

[Fos88] Nancy Foster. Citizens jam nuclear radio network. Bulletin of the Atomic Scientists, 44(9):21–26, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Foster:1995:CC

[Fos95] Gregory D. Foster. Clinton’s choice. Bulletin of the Atomic Scientists, 51(1):18–19, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sco95].

Fothergill:1960:BT

[Fot60] LeRoy D. Fothergill. The BW threat. Bulletin of the Atomic Scientists, 16(6):244–256, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fotadar:1989:LIA

[Fot89] Upinder Fotadar. Letter: India’s atoms for energy. Bulletin of the Atomic Scientists, 45(3):45, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spe89b].

Fotion:1991:BV

[Fot91a] Nicholas G. Fotion. Beyond Vietnam. Bulletin of the Atomic Scientists, 47(7):28, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HPR91].

Fotion:1991:GWC

[Fot91b] Nicholas G. Fotion. The Gulf War: Cleanly fought. Bulletin of the Atomic Scientists, 47(7):24–29, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fouere:1984:NGE

[Fou84] Erwan Fouéré. New global environment commission. Bulletin of the Atomic Scientists, 40(7):32–34, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fowler:1955:LED

[Fow55] Harold S. Fowler. Letter to the Editor: [Disarmament and the problem of peace]. Bulletin of the Atomic Scientists, 11(10):380– 381, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Szi55a].

Fowler:1958:LEC

[Fow58] Harold Fowler. Letter to the Editor: On the character of the Soviet economy. Bulletin of the Atomic Scientists, 14(2):96, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [dJ57b].

Fowler:1959:LES

[Fow59] Harold S. Fowler. Letter to the Editor: Survival kit. Bulletin of the Atomic Scientists, 15(1):49, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fox:1946:DWG

[Fox46] William T. R. Fox. Debate on world government or discussion of atomic energy control. Bulletin of the Atomic Scientists, 2(5–6): 22–23, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fox:1948:MRP

[Fox48] William T. R. Fox. “Middle-Run” planning: Atomic energy and international relations. Bulletin of the Atomic Scientists, 4(8):227–232, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fox:1952:LES

[Fox52] George A. Fox. Letter to the Editor: Senator McMahon’s speech. Bulletin of the Atomic Scientists, 8(2):65, February 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fox:1961:RWM

[Fox61] Irving K. Fox. Reason in water management. Bulletin of the Atomic Scientists, 17(10):418–422, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fox:1973:OLA

[Fox73] Sidney W. Fox. On the origin of life: The Apollo Program and amino acids. Bulletin of the Atomic Scientists, 29(10):46–51, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fox:1988:CDP

[Fox88] Thomas C. Fox. Catholics debate papal nuclear shift. Bulletin of the Atomic Scientists, 44(4):30–31, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:AE

[Foz61a] Paula Fozzy. Atomic energy. Bulletin of the Atomic Scientists, 17 (7):303–304, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:EWM

[Foz61b] Paula Fozzy. East and West meetings. Bulletin of the Atomic Scientists, 17(9):395–399, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:IN

[Foz61c] Paula Fozzy. International news. Bulletin of the Atomic Scientists, 17(7):302, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:NRC

[Foz61d] Paula Fozzy. News and reports: Civil defense developments. Bulletin of the Atomic Scientists, 17(8):347–351, October 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:NFS

[Foz61e] Paula Fozzy. News: Fallout shelters: Interest grows. Bulletin of the Atomic Scientists, 17(10):439–441, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:PM

[Foz61f] Paula Fozzy. Peace moves. Bulletin of the Atomic Scientists, 17 (7):302–303, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:PN

[Foz61g] Paula Fozzy. People in the news. Bulletin of the Atomic Scientists, 17(7):304, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:RE

[Foz61h] Paula Fozzy. Research–education. Bulletin of the Atomic Scientists, 17(7):304, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:RON

[Foz61i] Paula Fozzy. Russia opens new research areas. Bulletin of the Atomic Scientists, 17(7):300–301, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:SD

[Foz61j] Paula Fozzy. Space developments. Bulletin of the Atomic Scientists, 17(7):301–302, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1961:TBS

[Foz61k] Paula Fozzy. Test ban stymied. Bulletin of the Atomic Scientists, 17(7):299–300, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:NSC

[Foz62a] Paula Fozzy. New scientific consultant policy. Bulletin of the Atomic Scientists, 18(5):43–44, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:NAE

[Foz62b] Paula Fozzy. News: Atomic energy: New peaceful uses. Bulletin of the Atomic Scientists, 18(1):42–44, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:NLP

[Foz62c] Paula Fozzy. News: Lunar program underway. Bulletin of the Atomic Scientists, 18(2):37–40, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:NSE

[Foz62d] Paula Fozzy. News: Science and education: USSR. Bulletin of the Atomic Scientists, 18(4):43–48, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:NSF

[Foz62e] Paula Fozzy. News: Spring fallout increase expected. Bulletin of the Atomic Scientists, 18(3):43–44, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:RC

[Foz62f] Paula Fozzy. Research coordination. Bulletin of the Atomic Scientists, 18(3):29–31, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:SSR

[Foz62g] Paula Fozzy. Scientists’ social responsibility. Bulletin of the Atomic Scientists, 18(3):45–48, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:SC

[Foz62h] Paula Fozzy. Space cooperation. Bulletin of the Atomic Scientists, 18(5):45–46, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1962:TTB

[Foz62i] Paula Fozzy. Testing and the test ban. Bulletin of the Atomic Scientists, 18(6):46–47, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1963:MMV

[Foz63a] Paula Fozzy. Man on Mars and Venus. Bulletin of the Atomic Scientists, 19(5):24–25, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fozzy:1963:TVN

[Foz63b] Paula Fozzy. Two views of nuclear power: AEC, industry. Bulletin of the Atomic Scientists, 19(6):37–38, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Farinella:1989:LSV

[FP89] P. Farinella and V. V. Polcaro. Letter: Support Vanunu. Bulletin of the Atomic Scientists, 45(6):49, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mos88, CF87].

Florini:1990:BSS

[FP90a] Ann M. Florini and William C. Potter. Boon for Soviet space business. Bulletin of the Atomic Scientists, 46(9):31, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Florini:1990:GMC

[FP90b] Ann M. Florini and William C. Potter. Goodwill missions for castoff missiles. Bulletin of the Atomic Scientists, 46(9):25–31, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Florini:1990:RAT

[FP90c] Ann M. Florini and William C. Potter. Revamped Atlases and Titans. Bulletin of the Atomic Scientists, 46(9):28, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pow91a].

Findlay:2005:WW

[FP05] Trevor Findlay and Andreas Persbo. Watching the world. Bulletin of the Atomic Scientists, 61(2):58–63, March/April 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/2/58.

FitzGerald:2017:TGD

[FP17] Ben FitzGerald and Jacqueline Parziale. As technology goes democratic, nations lose military control. Bulletin of the Atomic Scientists, 73(2):102–107, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2017.
1288445.

Fermi:1976:MRH

[FR76] Enrico Fermi and Isadore I. Rabi. Minority report on the H-bomb: October 30, 1949. Bulletin of the Atomic Scientists, 32(10):58, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Herbert F. York’s book, The Advisors/Oppenheimer, Teller, and the superbomb.

Fairlie:1997:GCD

[FR97a] Ian Fairlie and Marvin Resnikoff. Global collective dose. Bulletin of the Atomic Scientists, 53(6):54, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fairlie:1997:NDT

[FR97b] Ian Fairlie and Marvin Resnikoff. No dose too low. Bulletin of the Atomic Scientists, 53(6):52–56, November/December 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Flynn:1998:BRM

[FR98] Michael Flynn and Linda Rothstein. Bulletin: The real mystery science theater. Bulletin of the Atomic Scientists, 54(6):9–10, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Franck:1947:STS

[Fra47] James Franck. The social task of the scientist … a scientist’s view of the social responsibilities of his profession — the honest dissemination of the truth. Bulletin of the Atomic Scientists, 3 (3):70, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Franck:1954:PFF

[Fra54] Peter G. Franck. Point Four — five years later. Bulletin of the Atomic Scientists, 10(6):218–223, 239, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1957:PVS

[Fra57] Philipp Frank. Part VIII: Science and philosophy: Philosophical uses of science. Bulletin of the Atomic Scientists, 13(4):125–130, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frampton:1958:RWR

[Fra58] Geoge T. Frampton. Radiation — whose responsibility? Bulletin of the Atomic Scientists, 14(10):421–425, December 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frampton:1961:BBE

[Fra61a] George T. Frampton. Books: Atomic Energy and Law. Inter-American Symposium, edited by Jaro Maydo. Bulletin of the Atomic Scientists, 17(10):437, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1961:AAP

[Fra61b] Jerome D. Frank. Atomic arms and pre-atomic man. Bulletin of the Atomic Scientists, 17(9):361–365, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1965:LES

[Fra65] Jerome D. Frank. Letter to the Editor: A statement on South Vietnam. Bulletin of the Atomic Scientists, 21(5):32–33, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1966:NWD

[Fra66] Lewis A. Frank. Nuclear weapons development in China. Bulletin of the Atomic Scientists, 22(1):12–15, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano66h] for a missing list of references.

Francoeur:1972:TFE

[Fra72] Robert T. Francoeur. Tomorrow’s fossils: Experimental embryology and endangered animal species. Bulletin of the Atomic Scientists, 28(2):11–14, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mag73].

Frasier:1974:IIM

[Fra74] E. Lewis Frasier. Improving an imperfect metric system. Bulletin of the Atomic Scientists, 30(2):9–12, 37–44, February 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wes74b].

Frank:1975:ICA

[Fra75] Forrest R. Frank. An international convention against nuclear theft. Bulletin of the Atomic Scientists, 31(10):51, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1976:IDS

[Fra76a] I. M. Frank. The importance of detant: a Soviet Nobel laureate calls on scientists to work for nuclear disarmament. Bulletin of the Atomic Scientists, 32(10):34–35, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1976:PAN

[Fra76b] Jerome D. Frank. Psychological aspects of the nuclear arms race. Bulletin of the Atomic Scientists, 32(4):22–24, April 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1979:PHN

[Fra79] Jerome M. Frank. Panel I: How a nuclear war can start … when fears take over. Bulletin of the Atomic Scientists, 35(4):25–27, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1980:PIY

[Fra80] Jerome D. Frank. Perspective: ‘If you win, you lose’. Bulletin of the Atomic Scientists, 36(8):1, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frank:1981:BRB

[Fra81] Jerome D. Frank. Book review: Landscapes of Fear, by Yi-fu Tuan. Bulletin of the Atomic Scientists, 37(6):44–45, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frankel:1982:BRB

[Fra82] Ben Frankel. Book review: The Politics of Uranium, by Norman Moss. Bulletin of the Atomic Scientists, 38(8):46, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frankel:1983:BRB

[Fra83] Ben Frankel. Book reviews: Controlling the Bomb: Nuclear Proliferation in the 1980s — A Twentieth Century Fund Report, by Lewis A. Dunn; Nuclear Power and Non-Proliferation: The Remaking of U.S. Policy, by Michael J. Brenner; Nuclear Poewr and Nonproliferation: An Interdisciplinary Perspective, by William C. Potter. Bulletin of the Atomic Scientists, 39(2):38–39, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frankel:1987:PMM

[Fra87] Sherman Frankel. Perspective: Moscow meeting — the glasnost menagerie. Bulletin of the Atomic Scientists, 43(4):8–12, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Franklin:1989:FFW

[Fra89] H. Bruce Franklin. Fatal fiction: a weapon to end all wars. Bulletin of the Atomic Scientists, 45(9):18–25, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [May90].

Frankel:1990:SAA

[Fra90a] Sherman Frankel. Stopping accidents after they’ve happened. Bulletin of the Atomic Scientists, 46(9):39–41, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano90-54].

Franklin:1990:LSJ

[Fra90b] H. Bruce Franklin. Letter: Saving Japan with the bomb. Bulletin of the Atomic Scientists, 46(4):52, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [May90].

Franklin:1991:BRZ

[Fra91] H. Bruce Franklin. Book review: The zen of Zinn: Declarations of Independence: Cross-Examining American Ideology, by Howard Zinn. Bulletin of the Atomic Scientists, 47(7):43–44, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Franz:2021:PIB

[Fra21] David R. Franz. Puzzling out the Iraqi biological weapons program. Bulletin of the Atomic Scientists, 77(4):166–171, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1941536.

Freistadt:1957:LEM

[Fre57] Hans Freistadt. Letter to the Editor: The magic of Marxism. Bulletin of the Atomic Scientists, 13(1):39–40, January 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pol56, Pol57a].

Freeman:1960:SPP

[Fre60] Orville L. Freeman. Science and party politics: The challenge of abundance. Bulletin of the Atomic Scientists, 16(9):372, 374, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1963:PNI

[Fre63] Orville L. Freeman. Planting new ideas: A report on a visit to Soviet farms. Bulletin of the Atomic Scientists, 19(10):44– 45, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fremlin:1967:CMB

[Fre67] J. H. Fremlin. Comments: Mind, brain, and humanist values. Bulletin of the Atomic Scientists, 23(2):26, February 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spe66].

Freedman:1968:UAP

[Fre68a] Daniel X. Freedman. The use and abuse of psychedelic drugs. Bulletin of the Atomic Scientists, 24(4):6–14, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fremlin:1968:MRB

[Fre68b] J. H. Fremlin. More reflections on big science. Bulletin of the Atomic Scientists, 24(7):33–34, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freedman:1971:TPE

[Fre71] S. David Freedman. Toward a policy of energy conservation. Bulletin of the Atomic Scientists, 27(8):8–12, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frei:1972:MAM

[Fre72a] E. H. Frei. Medical applications of magnetism: A new look at an archaic tool. Bulletin of the Atomic Scientists, 28(8):34–40, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Freund:1972:CMC

[Fre72b] Paul A. Freund. Can man control his biological evolution? — a symposium on genetic engineering: III. Xeroxing human beings. Bulletin of the Atomic Scientists, 28(10):18–20, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lut73].

Freeman:1973:NPE

[Fre73] S. David Freeman. Non-polluting energy: The energy joyride is over. Bulletin of the Atomic Scientists, 29(8):39–40, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1976:BEE

[Fre76a] S. David Freeman. Books: The energy establishment’s selffulfilling prophecies Beyond the Energy Crisis: A Global Perspective, by John Maddox. Bulletin of the Atomic Scientists, 32 (9):44–45, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1976:BFF

[Fre76b] S. David Freeman. Books: The facts and fictions of energy independence: Energy and World Politics, by Mason Willrich. Bulletin of the Atomic Scientists, 32(7):46–47, September 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fremlin:1976:LBP

[Fre76c] John H. Fremlin. Letter: Breeder protection. Bulletin of the Atomic Scientists, 32(7):3, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fremlin:1977:LNB

[Fre77a] J. H. Fremlin. Letter: No big problem. Bulletin of the Atomic Scientists, 33(6):5, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Maw77].

FAEC:1977:NFN

[Fre77b] French Atomic Energy Commission. A natural fossil nuclear reactor: How French scientists traced the existence of an ancient spontaneous chain reaction in a vast uranium deposit in West Africa. Bulletin of the Atomic Scientists, 33(2):40–41, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translation from French by Sarah Meselson.

Fremlin:1978:LOI

[Fre78] J. H. Fremlin. Letter: Obvious incentives. Bulletin of the Atomic Scientists, 34(4):3, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lor77].

Freedman:1981:CUK

[Fre81] Lawrence Freedman. Commentary from the United Kingdom: A critique of the END campaign. Bulletin of the Atomic Scientists, 37(10):38–42, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freedman:1983:NNF

[Fre83] Lawrence Freedman. Negotiations on nuclear forces. Bulletin of the Atomic Scientists, 39(10):22–28, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freedman:1984:BRBb

[Fre84a] Lawrence Freedman. Book review: The Reagan Strategic Defense Initiative: A Technical, Political, and Arms Control Assessment, by Sidney D. Drell, Philip J. Farley and David Holloway. Bulletin of the Atomic Scientists, 40(10):27–28, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freedman:1984:BRBa

[Fre84b] Lawrence J. Freedman. Book review: Living with Nuclear Weapons, by the Harvard Nuclear Study Group. Bulletin of the Atomic Scientists, 40(2):40–41, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freedman:1988:BRB

[Fre88] Lawrence Freedman. Book review: Policy versus the Law: The Reinterpretation of the ABM Treaty, by Raymond Garthoff. Bulletin of the Atomic Scientists, 44(3):51–52, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1991:C

[Fre91a] Karen Freeman. The chemicals. Bulletin of the Atomic Scientists, 47(10):34, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1991:GHD

[Fre91b] Karen Freeman. Grading the harm done. Bulletin of the Atomic Scientists, 47(10):38, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1991:UCW

[Fre91c] Karen Freeman. The unfought chemical war. Bulletin of the Atomic Scientists, 47(10):30–39, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1991:VSN

[Fre91d] Karen Freeman. A vast, secret network. Bulletin of the Atomic Scientists, 47(10):32, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1993:MBT

[Fre93a] Karen Freeman. The “man-break” tests. Bulletin of the Atomic Scientists, 49(2):40, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1993:VSA

[Fre93b] Karen Freeman. The VA’s sorry, the Army’s silent. Bulletin of the Atomic Scientists, 49(2):39–43, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Freeman:1995:BRB

[Fre95] Karen Freeman. Book review: Bioweapons horrors: Factories of Death: Japanese Biological Warfare 1932–45 and the American Cover Up, by Sheldon H. Harris. Bulletin of the Atomic Scientists, 51(1):63–64, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frelick:2002:NNB

[Fre02] Bill Frelick. Neglect is never benign. Bulletin of the Atomic Scientists, 58(6):26–35, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/6/26.

Freedman:2009:NTN

[Fre09] Lawrence Freedman. A new theory for nuclear disarmament. Bulletin of the Atomic Scientists, 65(4):14–30, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/4/14.

Friedwald:1948:ADC

[Fri48a] E. M. Friedwald. The atomic deadlock could be broken. Bulletin of the Atomic Scientists, 4(12):363–364, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mot48, Oli48, Ure48a] and reply [Fri48b].

Friedwald:1948:RTP

[Fri48b] E. M. Friedwald. Reply to the three professors. Bulletin of the Atomic Scientists, 4(12):366–367, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fri48a, Mot48, Oli48, Ure48a].

Friedman:1952:SFE

[Fri52] Georges Friedman. Some French experiences. Bulletin of the Atomic Scientists, 8(7):246, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Frisch:1954:FCD

[Fri54] O. R. Frisch. On the feasibility of coal-driven power stations. Bulletin of the Atomic Scientists, 10(6):224, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friendly:1956:PVS

[Fri56] Alfred Friendly. Part VI: Science, education, and communication: Scientists meet the press. Bulletin of the Atomic Scientists, 12 (9):338–340, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1959:PCN

[Fri59] O. R. Frisch. “Parity is not conserved” a new twist to physics? Bulletin of the Atomic Scientists, 15(3):139–143, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Universities Quarterly.

Frisch:1962:BBA

[Fri62] David Frisch. Books: Thinking About the Unthinkable, by Herman Kahn. Bulletin of the Atomic Scientists, 18(9):25–26, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedmann:1964:BBS

[Fri64] John Friedmann. Books: Cities and Space: The Future Use of Urban Space, edited by Lowdon Wingo. Bulletin of the Atomic Scientists, 20(5):30–32, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedman:1969:NTT

[Fri69] Edward Friedman. Now’s time to talk with China. Bulletin of the Atomic Scientists, 25(6):10–11, 35, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1970:SDB

[Fri70a] David H. Frisch. Scientists and the decision to bomb Japan. Bulletin of the Atomic Scientists, 26(6):107–115, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1970:BBF

[Fri70b] Rose Frisch. Books: The Future of the Future, by John McHale. Bulletin of the Atomic Scientists, 26(3):44, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1972:BBB

[Fri72] Rose E. Frisch. Books: The Biological Imperatives: Health, Politics and Human Survival, by Allan Chase. Bulletin of the Atomic Scientists, 28(9):46–47, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedman:1974:INO

[Fri74a] Todd Friedman. Israel’s nuclear option: The capability and the incentive. Bulletin of the Atomic Scientists, 30(7):30–36, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frisch:1974:STS

[Fri74b] Otto R. Frisch. Somebody turned the Sun on with a switch. Bulletin of the Atomic Scientists, 30(4):12–18, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedmann:1978:LMP

[Fri78] Herbert C. Friedmann. Letter: Man the player. Bulletin of the Atomic Scientists, 34(3):2, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Koe77].

Frisch:1979:HRU

[Fri79] David Frisch. Human rights and university contracts. Bulletin of the Atomic Scientists, 35(2):23–26, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HFMS79].

Fried:1982:CLN

[Fri82] John H. E. Fried. Commentary: Law and nuclear war. Bulletin of the Atomic Scientists, 38(6):67–68, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano82g].

Friedman:1984:CHR

[Fri84] Howard J. Friedman. Commentary: On ‘Human rights imperialism’. Bulletin of the Atomic Scientists, 40(3):62, March 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Von83].

Fridling:1988:LHP

[Fri88] Barry E. Fridling. Lasers highlight policy blindspots. Bulletin of the Atomic Scientists, 44(6):36, 38–39, July/August 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedman:1989:HGA

[Fri89] Mel Friedman. How Global Action got action. Bulletin of the Atomic Scientists, 45(6):17, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedman:1995:HM

[Fri95] Hideko Tamura Friedman. Hiroshima memories. Bulletin of the Atomic Scientists, 51(3):16–22, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedman:1997:PMP

[Fri97] John S. Friedman. Perspective: More power to thorium? Bulletin of the Atomic Scientists, 53(5):19–20, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rad97].

Friedman:2011:TMI

[Fri11] Sharon M. Friedman. Three Mile Island, Chernobyl, and Fukushima: an analysis of traditional and new media coverage of nuclear accidents and radiation. Bulletin of the Atomic Scientists, 67(5):55–65, September/October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/5/55.

Friedman:2020:HMO

[Fri20] Hideko Tamura Friedman. 1995: Hiroshima memories: One sunny day, a young girl learned about darkness. Bulletin of the Atomic Scientists, 76(6):397–403, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1847508.

Friederici:2021:GET

[Fri21] Peter Friederici. In Germany, the energy transition continues. Bulletin of the Atomic Scientists, 77(2):82–85, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1885851.

Frodin:1949:NSF

[Fro49] Reuben Frodin. National Science Foundation proposals: Background and commentary. Bulletin of the Atomic Scientists, 5(1):

21–24, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frutkin:1961:IPN

[Fru61] Arnold W. Frutkin. International programs of NASA. Bulletin of the Atomic Scientists, 17(5–6):229–232, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frutkin:1963:FSC

[Fru63] Arnold W. Frutkin. Forum: Space cooperation: International? Bulletin of the Atomic Scientists, 19(8):33–35, October 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fruchtbaum:1975:LC

[Fru75] Harold Fruchtbaum. Letters: A critique. Bulletin of the Atomic Scientists, 31(5):2, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75k].

Frye:1955:PUN

[Fry55] William R. Frye. Possession and use of nuclear weapons. Bulletin of the Atomic Scientists, 11(10):359–361, December 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1956:AHR

[Fry56a] William R. Frye. The atomic high road. Bulletin of the Atomic Scientists, 12(10):370–372, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1956:DD

[Fry56b] William R. Frye. The disarmament dilemma. Bulletin of the Atomic Scientists, 12(3):81–83, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1956:DBC

[Fry56c] William R. Frye. Disarmament: The bargain counter. Bulletin of the Atomic Scientists, 12(7):265–267, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1956:DTP

[Fry56d] William R. Frye. The disarmament turning point. Bulletin of the Atomic Scientists, 12(5):166–168, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1957:DUN

[Fry57a] William Frye. Disarmament in the United Nations: A new chapter? Bulletin of the Atomic Scientists, 13(3):91–94, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1957:DCB

[Fry57b] William R. Frye. Disarmament comes back to the General Assembly. Bulletin of the Atomic Scientists, 13(9):333–335, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Frye:1958:LER

[Fry58] William R. Frye. Letter to the Editor: Response to A. K. Mitra. Bulletin of the Atomic Scientists, 14(3):128, March 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mit58].

Frye:1965:FTN

[Fry65] Alton Frye. Forum: Three on nuclear policies: II. Space arms control. Bulletin of the Atomic Scientists, 21(4):30–33, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fry:1985:NAT

[Fry85] Greg Fry. Nuclear allergy: Toward a South Pacific nuclear-free zone. Bulletin of the Atomic Scientists, 41(6):16–20, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Friedman:1981:CDM

[FS81] Edward Friedman and Robert Schware. Commentary: Dispelling myths on carbon dioxide. Bulletin of the Atomic Scientists, 37 (9):55–56, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fingar:1984:STCa

[FS84] Thomas Fingar and Denis F. Simon. Science and technology in China: An overview. Bulletin of the Atomic Scientists, 40(8): 3s–5s [various between 24–25], October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fermi:1995:PBP

[FS95] Rachel Fermi and Esther Samra. Picturing the bomb: photographs from the secret world of the Manhattan Project. H. N. Abrams, New York, NY, USA, 1995. ISBN 0-8109-3735-2. 232 pp. LCCN QC773.A1 F47 1995. Introduction by Richard Rhodes.

Ferguson:2008:HBN

[FS08] Charles D. Ferguson and Michelle M. Smith. How not to build nuclear reactors. Bulletin of the Atomic Scientists, 64 (4):??, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/4/20.

Feiveson:1976:SIA

[FT76] Harold A. Feiveson and Theodore B. Taylor. Security implications of alternative fission futures. Bulletin of the Atomic Scientists, 32(10):14–18, 46–48, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feiveson:1976:PEW

[FTvHW76a] Harold A. Feiveson, Theodore B. Taylor, Frank von Hippel, and Robert H. Williams. The plutonium economy: Why we should wait and why we can wait. Bulletin of the Atomic Scientists, 32 (10):10–14, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feiveson:1976:SRP

[FTvHW76b] Harold A. Feiveson, Theodore B. Taylor, Frank von Hippel, and Robert H. Williams. [Sidebar:] the role of plutonium in nuclear power. Bulletin of the Atomic Scientists, 32(10):12–13, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fraley:1963:WWW

[FU63] Pierre C. Fraley and Earl Ubell. Writers writing & the written: Science writing: A growing profession. Bulletin of the Atomic Scientists, 19(10):19–22, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Fuchs:1970:PMD

[Fuc70] Sir Vivian Fuchs. Part 3: Management and development: Evolution of a venture in Antarctic science: Operation Tabarin and the British Antarctic Survey. Bulletin of the Atomic Scientists, 26 (10):75–80, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Fulbright:1969:MAM

[Ful69] J. W. Fulbright. Missiles and anti-missiles … six views: Foreign policy implications of the ABM debate. Bulletin of the Atomic Scientists, 25(6):20–23, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Furtwangler:1981:GN

[Fur81] Albert Furtwangler. Growing up nuclear. Bulletin of the Atomic Scientists, 37(1):44–48, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Feiveson:1985:RUS

[FUvH85] Harold A. Feiveson, Richard H. Ullman, and Frank von Hippel. Reducing U.S. and Soviet nuclear arsenals. Bulletin of the Atomic Scientists, 41(7):144–150, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lam85].

Fetter:1999:ADS

[FvH99] Steve Fetter and Frank von Hippel. After the dust settles. Bulletin of the Atomic Scientists, 55(6):42–45, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/6/42.

Feiveson:1986:BFW

[FvHA86] Harold A. Feiveson, Frank von Hippel, and David Albright. Breaking the fuel/weapons connection. Bulletin of the Atomic Scientists, 42(3):26–30, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Forbes:1964:LEW

[FW64] Allan Forbes and Arthur I. Waskow. Letter to the Editor: Without war. Bulletin of the Atomic Scientists, 20(6):36–38, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Was64].

Futter:2020:NWP

[FWCL20] Andrew Futter, Samuel I. Watson, Peter J. Chilton, and Richard J. Lilford. Nuclear war, public health, the COVID-19 epidemic: Lessons for prevention, preparation, mitigation, and education. Bulletin of the Atomic Scientists, 76(5):271–276, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1806592.

G:1965:CWN

[G.65] R. G. [Carrying the war to North Vietnam]. Bulletin of the Atomic Scientists, 21(4):2–3, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

G:1984:BNB

[G.84a] I. G. Book note: La Place de la Concorde Suisse, by John McPhee. Bulletin of the Atomic Scientists, 40(8):42, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

G:1984:BRB

[G.84b] L. G. Book review: World Armaments and Disarmament: SIPRI Yearbook 1984. Bulletin of the Atomic Scientists, 40(9):52–53, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geesaman:1973:FSP

[GA73] Donald P. Geesaman and Dean E. Abrahamson. Forensic science — a proposal. Bulletin of the Atomic Scientists, 29(3):17–24, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geesaman:1974:DFP

[GA74] Donald P. Geesaman and Dean E. Abrahamson. The dilemma of fission power. Bulletin of the Atomic Scientists, 30(9):37–41, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottemoeller:2008:ADN

[GA08] Rose Gottemoeller and Raymond Arnaudo. Agreeing to disagree on nuclear rights. Bulletin of the Atomic Scientists, 64(5):??, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/64/5/15.

Gabor:1973:BBC

[Gab73] Dennis Gabor. Books: The Collected Works of Leo Szilard, Scientific Papers, edited by Bernard T. Feld and Gertrud Weiss Szilard. Bulletin of the Atomic Scientists, 29(7):51–52, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gabriel:1986:LAS

[Gab86] Yousuf Gabriel. Letter: Appeal to scientists. Bulletin of the Atomic Scientists, 42(4):56, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gabelnick:2001:SAU

[Gab01] Tamar Gabelnick. Small arms: U.S., “Rogues” hold out. Bulletin of the Atomic Scientists, 57(5):18–19, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/5/18.

Gadler:1972:CAV

[Gad72] Steven J. Gadler. Communications: Another view of Monticello. Bulletin of the Atomic Scientists, 28(3):2, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sea71].

Gaddis:1989:ACH

[Gad89] John Lewis Gaddis. Arms control: Hanging tough paid off. Bulletin of the Atomic Scientists, 45(1):11–14, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kor89, Loe89].

Gaffney:1965:AEC

[Gaf65] Mason Gaffney. Applying economic controls. Bulletin of the Atomic Scientists, 21(6):20–25, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gaffney:1986:POR

[Gaf86] Frank J. Gaffney, Jr. Pentagon official responds to test ban challenge: [open letter to Edward J. Markey, US House of Representatives]. Bulletin of the Atomic Scientists, 42(4):11–12, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ber86c].

Gailar:1969:SWS

[Gai69] Joanne L. Gailar. Seven warning signals: A review of Soviet civil defense. Bulletin of the Atomic Scientists, 25(10):18–22, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gaines:1992:BRW

[Gai92] Linda Gaines. Book review: War games: Prisoner’s Dilemma, by William Poundstone. Bulletin of the Atomic Scientists, 48(8): 44–45, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gallini:1954:BBM

[Gal54] J. Gallini. Books: Microscopic Man, by Pierre Auger. Bulletin of the Atomic Scientists, 10(4):139–140, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gallois:1956:PIS

[Gal56] Pierre Gallois. Part II: Science and military strategy: The policy and strategy of air-nuclear weapons. Bulletin of the Atomic Scientists, 12(6):202–207, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gallois:1963:FTO

[Gal63] Pierre M. Gallois. Forum: The trap offered by de Gaulle. Bulletin of the Atomic Scientists, 19(8):24–27, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Sta63].

Galloway:1965:IRO

[Gal65] Eilene Galloway. International regulation of outer space activities. Bulletin of the Atomic Scientists, 21(2):36–39, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gall:1976:ABD

[Gal76] Norman Gall. Atoms for Brazil, dangers for all. Bulletin of the Atomic Scientists, 32(6):4–9, 41–48, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hay76].

Galal:1978:NGS

[Gal78a] Essam Eldin Galal. A new game: stabilization by accentuated destabilization. Bulletin of the Atomic Scientists, 34(2):39–44, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Galal:1978:P

[Gal78b] Essam Eldin Galal. A postscript. Bulletin of the Atomic Scientists, 34(2):43, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Galbraith:1981:CSE

[Gal81a] John Kenneth Galbraith. Commentary from Seattle: The economics of the arms race — and after. Bulletin of the Atomic Scientists, 37(6):13–16, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gallois:1981:CFM

[Gal81b] Pierre M. Gallois. Commentary: French military politics. Bulletin of the Atomic Scientists, 37(7):21–25, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kol81].

Gallagher:1986:BRB

[Gal86] Matthew Gallagher. Book review: Detente and Confrontation: American–Soviet Relations from Nixon to Reagan, by Raymond L. Garthoff. Bulletin of the Atomic Scientists, 42(8):49–51, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gale:1990:CAS

[Gal90a] Robert Peter Gale. Chernobyl: Answers slipping away. Bulletin of the Atomic Scientists, 46(7):19–23, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gale:1990:RHC

[Gal90b] Robert Peter Gale. Radiation and health: Chernobyl: Answers slipping away. Bulletin of the Atomic Scientists, 46(7):19–23, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gallagher:1991:RNP

[Gal91] Michael C. Gallagher. Report: Nuclear power: Hong Kong fears Chinese Chernobyl. Bulletin of the Atomic Scientists, 47(8):9– 11, 46, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gallagher:2013:NPM

[Gal13] Carole Gallagher. Nuclear photography: Making the invisible visible. Bulletin of the Atomic Scientists, 69(6):42–46, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/69/6/42.

Gamova:1994:MTS

[Gam94] Svetlana Gamova. Moldova:a three-say split. Bulletin of the Atomic Scientists, 50(1):41–43, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gantt:1952:RSS

[Gan52] W. Horsley Gantt. Russian science symposium: VII: Bolshevik principles and Russian physiology. Bulletin of the Atomic Scientists, 8(6):183–189, 206, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gandhi:1972:SRW

[Gan72] Indira Gandhi. A special report: What happened at Stockholm. V. What they said — Indira Gandhi, India. Bulletin of the Atomic Scientists, 28(7):35–, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gandhi:1976:DRP

[Gan76] Indira Gandhi. Different roads to peace and justice. Bulletin of the Atomic Scientists, 32(4):13–15, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sma76b].

Ganguly:1982:BRB

[Gan82] Sumit Ganguly. Book review: The Global Politics of Arms Sales, by Andrew J. Pierre. Bulletin of the Atomic Scientists, 38(9):38– 39, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:1983:BNB

[Gan83a] Sumit Ganguly. Book note: The Nuclear Future, by Michael Mandelbaum. Bulletin of the Atomic Scientists, 39(8):38, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:1983:BRBa

[Gan83b] Sumit Ganguly. Book review: Nuclear India: A Technological Assessment, by G. G. Mirchandani and P. K. S. Namboodri; Nuclear Myths and Realities: India’s Dilemma. Bulletin of the Atomic Scientists, 39(3):42–43, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:1983:BRBb

[Gan83c] Sumit Ganguly. Book review: Nuclear Power in India, by David Hart. Bulletin of the Atomic Scientists, 39(10):44, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:1983:SP

[Gan83d] Sumit Ganguly. Student Pugwash. Bulletin of the Atomic Scientists, 39(9):50–51, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:1983:WIJ

[Gan83e] Sumit Ganguly. Why India joined the nuclear club. Bulletin of the Atomic Scientists, 39(4):30–33, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:1984:BNB

[Gan84] Sumit Ganguly. Book note: Mimicking Sisyphus: America’s Countervailing Nuclear Strategy, by Louis René Beres. Bulletin of the Atomic Scientists, 40(3):48, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ganguly:2003:KCM

[Gan03] Sumit Ganguly. Kashmir, caught in the middle. Bulletin of the Atomic Scientists, 59(4):25–28, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/4/25.

Gao:2017:UPD

[Gao17] Wanglai Gao. Unearthing poison: Disposal of abandoned chemical weapons in China. Bulletin of the Atomic Scientists, 73 (6):404–410, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1388678.

Gapes:1986:LLP

[Gap86] Mike Gapes. Letter: Labour Party’s nuclear policy. Bulletin of the Atomic Scientists, 42(9):56, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gro86] and comment [Ken87c].

Gapes:1990:LLR

[Gap90] Mike Gapes. Letter: Labor’s real defense policy. Bulletin of the Atomic Scientists, 46(1):45, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mos89].

Garrison:1954:ORR

[Gar54] Lloyd K. Garrison. Oppenheimer requests review. Bulletin of the Atomic Scientists, 10(6):251–254, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter sent to General Nichols by the author, who is the counsel for J. Robert Oppenheimer.

Gardner:1960:OP

[Gar60] Trevor Gardner. Organizing for peace. Bulletin of the Atomic Scientists, 16(7):297–300, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garcia:1966:OSR

[Gar66] Rolando V. Garcia. Organizing scientific research. Bulletin of the Atomic Scientists, 22(7):12–15, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1967:MOM

[Gar67] R. L. Garwin. MIRV and the offensive missile race. Bulletin of the Atomic Scientists, 23(10):21, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garfield:1968:CAS

[Gar68] Eugene Garfield. Chemical Abstracts Service annual report to NSF. Bulletin of the Atomic Scientists, 24(6):43–44, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1972:BBP

[Gar72] Richard L. Garwin. Books: Strategic Power and National Security, by J. I. Coffey. Bulletin of the Atomic Scientists, 28(7):59– 60, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garson:1974:LPT

[Gar74] Paul Richard Garson. Letter: A perception test. Bulletin of the Atomic Scientists, 30(7):3, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor74].

Gardner:1977:GFC

[Gar77a] John Gardner. The gang of four and Chinese science. Bulletin of the Atomic Scientists, 33(7):24–30, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garrett:1977:DMB

[Gar77b] Stephen A. Garrett. Détente and the military balance. Bulletin of the Atomic Scientists, 33(4):10–20, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1977:LPL

[Gar77c] Richard L. Garwin. Letter: The promise of laser isotope separation. Bulletin of the Atomic Scientists, 33(8):8–9, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dic77].

Garfield:1978:WCN

[Gar78a] Eugene Garfield. What this country needs is a free phone call. Bulletin of the Atomic Scientists, 34(2):56–59, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1978:CPB

[Gar78b] R. L. Garwin. Charged-particle beam weapons? Bulletin of the Atomic Scientists, 34(8):24–27, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1981:SRW

[Gar81] Richard L. Garwin. Special report: Are we on the verge of an arms-race in space? Bulletin of the Atomic Scientists, 37(5): 48–53, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gardiner:1983:BRB

[Gar83a] Richard Gardiner. Book review: Radiation Protection: A Guide for Scientists and Physicians, second edition, by Jacob Shapiro. Bulletin of the Atomic Scientists, 39(1):36–37, January 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1983:WPW

[Gar83b] Richard L. Garwin. Who proposes, who disposes, who pays? Bulletin of the Atomic Scientists, 39(6):9–11, June/July 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SG83].

Garthoff:1984:AAC

[Gar84a] Raymond L. Garthoff. ASAT arms control: still possible. Bulletin of the Atomic Scientists, 40(7):29–31, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garthoff:1984:WSG

[Gar84b] Raymond L. Garthoff. From Washington: The ‘spending gap’. Bulletin of the Atomic Scientists, 40(5):5–6, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garthoff:1984:PIT

[Gar84c] Raymond L. Garthoff. Postmortem on INF talks. Bulletin of the Atomic Scientists, 40(10):7–10, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garrison:1985:LJF

[Gar85] Lloyd W. Garrison. Letter: Journalistic freedom and responsibility. Bulletin of the Atomic Scientists, 41(5):56, May 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1986:LSD

[Gar86] Richard L. Garwin. Letter: SDI doubletalk. Bulletin of the Atomic Scientists, 42(6):61, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DM86].

Garthoff:1987:RSD

[Gar87] Raymond L. Garthoff. Refocusing the SDI debate. Bulletin of the Atomic Scientists, 43(7):44–50, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bal87].

Gardner:1988:UCN

[Gar88a] David Gardner. The University of California: “No further action is appropriate”. Bulletin of the Atomic Scientists, 44(6):12, July/ August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gareyev:1988:RSM

[Gar88b] Makhmut Gareyev. The revised Soviet military doctrine. Bulletin of the Atomic Scientists, 44(10):30–34, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lei89a].

Garthoff:1988:DKB

[Gar88c] Raymond L. Garthoff. Did Khrushchev bluff in Cuba? No. Bulletin of the Atomic Scientists, 44(6):40–43, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Leb88b, Leb88a].

Garwin:1988:BRW

[Gar88d] Richard Garwin. A blueprint for radical weapons cuts. Bulletin of the Atomic Scientists, 44(2):10–13, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garfinkle:1989:ACC

[Gar89] Adam M. Garfinkle. Arms control criteria: The world has changed. Bulletin of the Atomic Scientists, 45(4):44, May 1989.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garthoff:1991:RSS

[Gar91] Raymond L. Garthoff. Report: Soviet snafu: Case of the wandering radar. Bulletin of the Atomic Scientists, 47(6):7–9, July/ August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1992:DNN

[Gar92] Richard L. Garwin. Don’t neglect new weapons plant. Bulletin of the Atomic Scientists, 48(4):17–18, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1993:PAD

[Gar93] Richard L. Garwin. Perspective: Atoms do not age. Bulletin of the Atomic Scientists, 49(8):10–11, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garthoff:1997:BRP

[Gar97a] Raymond L. Garthoff. Book review: Provocative hypotheses: We Now Know: Rethinking Cold War History, by John Lewis Gaddis. Bulletin of the Atomic Scientists, 53(6):58–60, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1997:LTT

[Gar97b] Richard Garwin. Letter: Two tracks are better than one. Bulletin of the Atomic Scientists, 53(2):3, 60, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vHJ96, Lym96].

Garwin:1997:PSD

[Gar97c] Richard L. Garwin. Perspective: Stewardship: Don’t claim too much or too little. Bulletin of the Atomic Scientists, 53(3):21– 24, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1998:RRW

[Gar98] Richard L. Garwin. The Rumsfeld report: what we did. Bulletin of the Atomic Scientists, 54(6):40–45, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garthoff:2000:PR

[Gar00a] Raymond L. Garthoff. Polyakov’s run. Bulletin of the Atomic Scientists, 56(5):37–40, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/5/37.

Garwin:2000:WP

[Gar00b] Richard L. Garwin. The wrong plan. Bulletin of the Atomic Scientists, 56(2):36–41, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/2/36.

Garwin:2008:EIM

[Gar08] Richard L. Garwin. Evaluating Iran’s missile threat. Bulletin of the Atomic Scientists, 64(2):40–43, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/2/40.

Gascoyne:1993:SSS

[Gas93a] Stephen Gascoyne. Slipcovering a Superfund site. Bulletin of the Atomic Scientists, 49(7):33–37, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gascoyne:1993:SBI

[Gas93b] Stephen Gascoyne. SQI — the burning issue. Bulletin of the Atomic Scientists, 49(7):34–35, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gat:1972:WRR

[Gat72] Joel R. Gat. Water resources research in Israel. Bulletin of the Atomic Scientists, 28(8):24–28, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Gaulkin:2024:CVW

[Gau24] Thomas Gaulkin. The campaign volunteer who used AI to help swing Pakistan’s elections: Interview with Jibran Ilyas. Bulletin of the Atomic Scientists, 80(5):275–280, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2388458.

Gavin:1951:TUA

[Gav51] Major General James M. Gavin. The tactical user of the atomic bomb. Bulletin of the Atomic Scientists, 7(2):46–47, 50, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gayler:1983:BND

[Gay83] Noel Gayler. Breaking the nuclear deadlock: A proposal for deep cuts. Bulletin of the Atomic Scientists, 39(10):46–47, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gay:1998:BWH

[Gay98] Corey Gay. Bulletin: So, whatever happened to checking with another source? Bulletin of the Atomic Scientists, 54(4):12, July/ August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gollub:1972:CT

[GB72] Jerry P. Gollub and R. Stephen Berry. Communications: Thermodynamics. Bulletin of the Atomic Scientists, 28(9):3–4, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber72c].

Glaser:1973:NDS

[GB73] Peter E. Glaser and James C. Burke. New directions for solar energy. Bulletin of the Atomic Scientists, 29(8):40–42, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1977:WWM

[GB77] Robert C. Gray and Robert J. Bresler. Why weapons make poor bargaining chips. Bulletin of the Atomic Scientists, 33(7):8–9, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldanski:1988:SSS

[GB88] Vitali I. Goldanski and Vanda Beletskaya. Struggles of a Soviet scientist. Bulletin of the Atomic Scientists, 44(8):33–35, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translated from Russian to English by Kurt Campbell.

Gale:2011:IUB

[GB11] Robert Peter Gale and Alexander Baranov. If the unlikely becomes likely: Medical response to nuclear accidents. Bulletin of the Atomic Scientists, 67(2):10–18, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/67/2/10.

Gueritz:1981:CNB

[GBG^{+}81] Edward F. Gueritz, Georges Buis, Wolf Graf von Baudissin, Michael Harbottle, Pierre Lellouche, Philipp Sonntag, Gerard C. Berkhof, Inga Thorsson, and R. J. H. Kruisinga. Commentary: The neutron bomb: Reactions from Europe. Bulletin of the Atomic Scientists, 37(8):7–13, 59–60, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greenburg:1963:LEO

[GC63] David F. Greenburg and Tucker Carrington. Letter to the Editor: The Old Adam. Bulletin of the Atomic Scientists, 19(6):30–31, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vL63].

Glackin:1977:LLP

[GC77] James Glackin and Barry M. Casper. Letter: Laser plutonium separation. Bulletin of the Atomic Scientists, 33(4):3, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cas77b].

Grobstein:1983:CTP

[GC83] Clifford Grobstein and Liebe Cavalieri. Commentary: On ‘Twin Perils: Nuclear Science and Genetic Engineering’. Bulletin of the Atomic Scientists, 39(4):48, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cav82].

Gross:2016:PEC

[GCV16] Michael L. Gross, Daphna Canetti, and Dana R. Vashdi. The psychological effects of cyber terrorism. Bulletin of the Atomic Scientists, 72(5):284–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grimsson:1985:IGF

[GD85] Olafur Grimsson and Nicholas Dunlop. Indira Gandhi and the Five Continent Initiative. Bulletin of the Atomic Scientists, 41 (1):46, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gromyko:2004:RJR

[GDH^{+}04] A. Gromyko, F. Dyson, K. C. Hart, P. H. Nitze, A. M. Smith, S. L. Rearden, M. Moore, J. Reston, R. Aron, and A. Mill. Remembering J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 60(2):37–??, March/April 2004. CODEN BASIAP. ISSN 0096-3402.

Geary:1989:NTS

[Gea89] Riley R. Geary. Nevada Test Site’s dirty little secrets. Bulletin of the Atomic Scientists, 45(3):35–36, 38, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [EG89, Han89].

Geertz:1966:ICC

[Gee66] Clifford Geertz. The impact of the concept of culture and the concept of man. Bulletin of the Atomic Scientists, 22(4):2–8, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geesaman:1971:PED

[Gee71] Donald P. Geesaman. Plutonium and the energy decision. Bulletin of the Atomic Scientists, 27(7):33–36, September 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geiger:1981:CSF

[Gei81] H. Jack Geiger. Commentary from San Francisco: The illusion of ‘survival’. Bulletin of the Atomic Scientists, 37(6):16–20, June/ July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geist:2016:ATL

[Gei16a] Edward Moore Geist. It’s already too late to stop the AI arms race — we must manage it instead. Bulletin of the Atomic Scientists, 72(5):318–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geist:2016:WRU

[Gei16b] Edward Moore Geist. Would Russia’s undersea “doomsday drone” carry a cobalt bomb? Bulletin of the Atomic Scientists, 72(4):238–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geist:2017:APT

[Gei17] Edward Moore Geist. (Automated) planning for tomorrow: Will artificial intelligence get smarter? Bulletin of the Atomic Scientists, 73(2):80–85, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1288435.

Gellner:1990:LIV

[Gel90] Charles R. Gellner. Letter: Israel’s value as ally disputed. Bulletin of the Atomic Scientists, 46(9):46–47, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei90, Sol90].

GACUSAEC:1959:PPR

[Gen59] General Advisory Committee to the U.S.Atomic Energy Commission. Problems presented by radioactive fallout. Bulletin of the Atomic Scientists, 15(6):258–259, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gendlin:1971:PPP

[Gen71] Frances Gendlin. The Palisades protest: A pattern of citizen intervention. Bulletin of the Atomic Scientists, 27(9):53–56, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gendin:1972:SRW

[Gen72] Frances Gendin. A special report: What happened at Stockholm: III. Voices from the gallery. Bulletin of the Atomic Scientists, 28 (7):26–29, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Genuth:1984:BNB

[Gen84a] Joel Genuth. Book note: Scientific Temperaments: Three Lives in Contemporary Science, by Philip J. Hilts. Bulletin of the Atomic Scientists, 40(6):39–40, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Genuth:1984:BRBa

[Gen84b] Joel Genuth. Book review: America’s Unwritten Constitution: Science, Religion, and Political Responsibility, by Don K. Price.

Bulletin of the Atomic Scientists, 40(3):43–45, March 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Genuth:1984:BRBb

[Gen84c] Joel Genuth. Book review: The Discoverers: A History of Man’s Search to Know His World and Himself, by Daniel J. Boorstin. Bulletin of the Atomic Scientists, 40(7):40–42, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Genuth:1986:BRB

[Gen86] Joel Genuth. Book reviews: Radiant Science, Dark Politics: A Memoir of the Nuclear Age, by Martin D. Kamen; The Education of a College President, by James R. Killian, Jr. Bulletin of the Atomic Scientists, 42(1):51–52, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gerard:1949:SR

[Ger49] R. W. Gerard. The scientific reserve. Bulletin of the Atomic Scientists, 5(10):276–280, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gerard:1958:LER

[Ger58] Ralph W. Gerard. Letter to the Editor: The Russian Sputnik. Bulletin of the Atomic Scientists, 14(9):401, November 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gerber:1967:CAV

[Ger67] Alex Gerber. Comments: After victory, what? Bulletin of the Atomic Scientists, 23(3):27–28, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gerasimov:1972:CLD

[Ger72] Gennady Gerasimov. Communications: On living dangerously. Bulletin of the Atomic Scientists, 28(5):4, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab72g].

Gerrish:1975:LTM

[Ger75] Laina Gerrish. Letters: Too much. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Gerjuoy:1982:EIR

[Ger82] Edward Gerjuoy. Embargo on ideas: the Reagan administration. Bulletin of the Atomic Scientists, 38(9):31–37, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gerson:1984:CNB

[Ger84] Joseph Gerson. Commentary: Nuclear blackmail. Bulletin of the Atomic Scientists, 40(5):57–58, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sha85a].

Gerson:1996:R

[Ger96] Joseph Gerson. “I refuse”. Bulletin of the Atomic Scientists, 52 (4):26, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gerami:2013:ACW

[Ger13] Nima Gerami. Attracting a crowd: What societal verification means for arms control : The US response. Bulletin of the Atomic Scientists, 69(3):14–18, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/3/14.

Gerstein:2017:GGA

[Ger17] Daniel M. Gerstein. Glaring gaps: America needs a biodefense upgrade. Bulletin of the Atomic Scientists, 73(2):86–91, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1288438.

Gerstein:2020:AUG

[Ger20] Daniel M. Gerstein. Assessing the US government response to the coronavirus. Bulletin of the Atomic Scientists, 76(4):166– 174, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1778356.

Gerstein:2021:MPR

[Ger21] Daniel M. Gerstein. Memo to the President: Reimaging public health preparedness and response. Bulletin of the Atomic Scientists, 77(1):41–42, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1859864.

Gevorkian:1993:KTS

[Gev93] Natalia Gevorkian. The KGB: “They still need us”. Bulletin of the Atomic Scientists, 49(1):36–38, January/February 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geyer:1994:GOW

[Gey94] Georgie Anne Geyer. Guest opinion: Walking too softly. Bulletin of the Atomic Scientists, 50(4):4, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sco94b].

Gerber:1964:LEJ

[GF64] Alex Gerber and Frederick Forscher. Letter to the Editor: On de Jouvenel. Bulletin of the Atomic Scientists, 20(4):32, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [dJ63].

Goldstein:1973:CAQ

[GF73] Paul Goldstein and Robert Ford. On the control of air quality: Why the laws don’t work. Bulletin of the Atomic Scientists, 29 (6):31–34, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1979:CME

[GF79] I. I. Glass and Bernard T. Feld. Commentary: Muddled editorials? Bulletin of the Atomic Scientists, 35(4):55–56, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottlieb:1963:LEC

[GG63] Sanford Gottlieb and Bentley Glass. Letter to the Editor: Congress and disarmament. Bulletin of the Atomic Scientists, 19 (10):40–41, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cla63].

Gandhi:1987:NDD

[GG87] Rajiv Gandhi and Mikhail Gorbachev. New Delhi declaration. Bulletin of the Atomic Scientists, 43(2):37, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldston:2011:ICF

[GG11a] Robert J. Goldston and Alexander Glaser. Inertial confinement fusion energy R&D and nuclear proliferation: The need for direct and transparent review. Bulletin of the Atomic Scientists, 67 (3):59–66, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/3/59.

Granoff:2011:IHL

[GG11b] Dean Granoff and Jonathan Granoff. International humanitarian law and nuclear weapons: Irreconcilable differences. Bulletin of the Atomic Scientists, 67(6):53–62, November/December 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/6/53.

Gifford:2016:LUI

[GG16] Eva Gifford and Robert Gifford. The largely unacknowledged impact of climate change on mental health. Bulletin of the Atomic Scientists, 72(5):292–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greb:1983:LHT

[GH83] G. Allen Greb and Warren Heckrotte. The long history: the test ban debate. Bulletin of the Atomic Scientists, 39(7):36–42, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mar83d].

Ganguly:1985:SP

[GH85] Sumit Ganguly and David Hart. Student Pugwash 1985. Bulletin of the Atomic Scientists, 41(11):58, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1994:LWA

[GH94] Stanley Goldberg and David Hawkins. Letters: Wrong about Bohr, too. Bulletin of the Atomic Scientists, 50(5):3, 59–60, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ghoshroy:2011:CST

[Gho11] Subrata Ghoshroy. Coming not so soon to a theater near you: Laser weapons for missile defense. Bulletin of the Atomic Scientists, 67(6):34–43, November/December 2011. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/6/34.

Ghosh:2015:BPR

[Gho15a] Arunabha Ghosh. The big push for renewable energy in India: What will drive it? Bulletin of the Atomic Scientists, 71(4): 31–42, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ghoshroy:2015:XBW

[Gho15b] Subrata Ghoshroy. The X-37B: Backdoor weaponization of space? Bulletin of the Atomic Scientists, 71(3):19–29, May 2015. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gibbons:1982:CDM

[Gib82] John H. Gibbons. [Comment on Decision-making in a democracy]. Bulletin of the Atomic Scientists, 38(5):49–50, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Scl82].

Gibson:1991:CAS

[Gib91a] Daniel Gibson. Can alchemy solve the nuclear waste problem? Bulletin of the Atomic Scientists, 47(6):12–17, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PG91].

Gibson:1991:TRP

[Gib91b] Daniel Gibson. Transmutation-related projects. Bulletin of the Atomic Scientists, 47(6):16, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Giddings:1973:WPH

[Gid73] J. Calvin Giddings. World population, human disaster and the nuclear holocaust. Bulletin of the Atomic Scientists, 29(7):21– 24, 45–50, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gifford:2013:DMH

[Gif13] Robert Gifford. Dragons, mules, and honeybees: Barriers, carriers, and unwitting enablers of climate change action. Bulletin of the Atomic Scientists, 69(4):41–48, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/69/4/41.

Gigon:1960:FD

[Gig60] Fernand Gigon. Formula for death. Bulletin of the Atomic Scientists, 16(9):359, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). A sad note on the fate of decorated pilot Major Claude R. Eatherly, who flew over Hiroshima 15 minutes after the atomic bomb was dropped there, and three days later, piloted the plane that dropped an atomic bomb on Nagasaki. Eatherly’s post-war behavior and symptoms described there would likely be recognized now as post-traumatic stress disorder (PTSD).

Gilson:1950:NF

[Gil50] Etienne Gilson. Neutrality for France? Bulletin of the Atomic Scientists, 6(7):203–205, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilbert:1955:VRI

[Gil55] Horace N. Gilbert. Visa regulations and the international exchange of students. Bulletin of the Atomic Scientists, 11(10): 371–373, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gill:1963:FNP

[Gil63] R. Rockingham Gill. Forum: Nuclear power. Bulletin of the Atomic Scientists, 19(7):32, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilden:1965:LES

[Gil65] K. B. Gilden. Letter to the Editor: Science and the shabby curate. Bulletin of the Atomic Scientists, 21(9):24, November 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim65].

Gill:1968:DSS

[Gil68] Richard Rockingham Gill. Decisionmaking in Soviet science policy. Bulletin of the Atomic Scientists, 24(4):15–19, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilgore:1969:PTCb

[Gil69a] Sheldon G. Gilgore. Pharmaceutical testing: Company view: A further comment. Bulletin of the Atomic Scientists, 25(6):16–17, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilgore:1969:PTCa

[Gil69b] Sheldon G. Gilgore. Pharmaceutical testing: Company view: The way it really is. Bulletin of the Atomic Scientists, 25(6):12, 18–19, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gillis:1972:SED

[Gil72] Joseph Gillis. Science education in a developing country: Reforming the curriculum. Bulletin of the Atomic Scientists, 28(8): 41–44, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Gilbert:1977:LS

[Gil77] Joseph A. Gilbert. Letter: Any suggestions? Bulletin of the Atomic Scientists, 33(7):6, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilinsky:1980:ITM

[Gil80] V. Gilinsky. The impact of Three Mile Island. Bulletin of the Atomic Scientists, 36(1):18–20, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilinsky:1983:BRB

[Gil83] Victor Gilinsky. Book review: Nuclear Power: Both Sides: The Best Arguments For and Against the Most Controversial Technology, edited by Michio Kaku and Jennifer Trainer. Bulletin of the Atomic Scientists, 39(5):40–41, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilbertson:1984:H

[Gil84] Philip N. Gilbertson. Humanities. Bulletin of the Atomic Scientists, 40(10):13s–15s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilkey:1985:BRB

[Gil85] Langdon Gilkey. Book review: The Limits of Science, by P. B. Medawar. Bulletin of the Atomic Scientists, 41(9):50–51, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gilinsky:1998:LNK

[Gil98] Victor Gilinsky. Letter: North Korea’s blackmail. Bulletin of the Atomic Scientists, 54(2):3, 68, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sig98].

Gilinsky:2018:CFH

[Gil18] Victor Gilinsky. A containment failure: How American nuclear regulators undercut power plant safety from the beginning. Bulletin of the Atomic Scientists, 74(3):170–176, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1461975.

Ginsburg:1954:BBN

[Gin54] Norton S. Ginsburg. Books: Japan’s Natural Resources, by Edward A. Ackerman. Bulletin of the Atomic Scientists, 10(3):97– 98, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ginsburg:1976:LPL

[Gin76] Norton Ginsburg. Letter: Playing with lives. Bulletin of the Atomic Scientists, 32(9):6–7, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Girdner:1975:LNA

[Gir75] William Girdner. Letters: No alternative? Bulletin of the Atomic Scientists, 31(6):4–5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Gitermann:1954:SHS

[Git54] V. Gitermann. The study of history in the Soviet Union. Bulletin of the Atomic Scientists, 10(6):206–208, 240, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gittings:1969:PCR

[Git69] John Gittings. The prospects of the Cultural Revolution in 1969. Bulletin of the Atomic Scientists, 25(2):23–28, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Gitterman:1974:FER

[Git74] Moisei Gitterman. On the freedom of emigration: A Russian emigré to Israel pleads for his colleagues. Bulletin of the Atomic Scientists, 30(2):36–367, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1967:CEG

[GJ67] Lazer Goldberg and Nehemiah Jordan. Comments: An exchange — Goldberg and Jordan. Bulletin of the Atomic Scientists, 23 (4):34, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jor67].

Greb:1984:WTP

[GJ84] G. Allen Greb and Gerald W. Johnson. The weapons tutorial: [part 7:] a history of strategic arms limitations. Bulletin of the Atomic Scientists, 40(1):30–37, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Good:1980:CSS

[GJS80] Norman E. Good, Thomas B. Johansson, and Peter Steen. Commentary: On ‘solar Sweden’. Bulletin of the Atomic Scientists, 36(3):59–60, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [JS79a].

Graybeal:1985:SNA

[GK85] Sidney N. Graybeal and Michael Krepon. SCC: neglected arms control tool. Bulletin of the Atomic Scientists, 41(10):30–33, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Graybeal:1987:LSI

[GK87] Sidney N. Graybeal and Michael Krepon. The limitations of onsite inspection. Bulletin of the Atomic Scientists, 43(10):22–26, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DeV88].

Graybeal:1994:PSS

[GK94] Sidney N. Graybeal and Michael Krepon. Perspective: It’s not son of Star Wars. Bulletin of the Atomic Scientists, 50(2):16–17, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [TKG94].

Goodby:2021:AAC

[GK21] James E. Goodby and David A. Koplow. An ambitious arms control agenda requires a new organization equal to the task. Bulletin of the Atomic Scientists, 77(1):5–10, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1859858.

Gerson:1948:ALP

[GL48] M. Gerson and M. L. Lesser. AEC loyalty procedures and civil rights. Bulletin of the Atomic Scientists, 4(7):197–198, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Geiger:1990:FLS

[GL90] H. Jack Geiger and Ken Lichtenstein. A first look at the studies. Bulletin of the Atomic Scientists, 46(7):26, September 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Golovkov:1995:TG

[GL95] Aleksandr Golovkov and Sergei Leskov. Top guns. Bulletin of the Atomic Scientists, 51(3):51–54, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1956:AFT

[Gla56a] Bentley Glass. Academic freedom and tenure in the quest for national security. Bulletin of the Atomic Scientists, 12(6):221– 222, 226, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1956:HAR

[Gla56b] Bentley Glass. The hazards of atomic radiations to man — British and American reports. Bulletin of the Atomic Scientists, 12(8):312–317, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1958:LES

[Gla58] Bentley Glass. Liberal education in a scientific age. Bulletin of the Atomic Scientists, 14(9):346–353, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1960:GEB

[Gla60] Bentley Glass. Guest editorial: Biological and chemical warfare: An international symposium. Bulletin of the Atomic Scientists, 16(6):226–227, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glagolev:1962:DP

[Gla62a] Igor Glagolev. Is disarmament practical? Bulletin of the Atomic Scientists, 18(7):11–13, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1962:ICB

[Gla62b] Bentley Glass. Information crisis in biology. Bulletin of the Atomic Scientists, 18(8):6–12, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1962:SP

[Gla62c] Bentley Glass. Scientists in politics. Bulletin of the Atomic Scientists, 18(5):2–7, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response and reply [RGR63].

Glass:1963:LEN

[Gla63] Bentley Glass. Letter to the Editor: No more war. Bulletin of the Atomic Scientists, 19(10):41, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RGR63].

Glass:1968:CU

[Gla68] Bentley Glass. Crisis in the universities. Bulletin of the Atomic Scientists, 24(8):2–3, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glackin:1975:HSP

[Gla75a] James J. Glackin. How secrecy played executioner: Would the Rosenbergs have been sentenced to die if the prosecution had told all it knew? Bulletin of the Atomic Scientists, 31(6):14–16, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glackin:1975:LIE

[Gla75b] James J. Glackin. Letter: The impact of energy. Bulletin of the Atomic Scientists, 31(9):3–4, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glackin:1975:LLS

[Gla75c] James J. Glackin. Letters: A lesser solution. Bulletin of the Atomic Scientists, 31(2):2, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glackin:1975:LPE

[Gla75d] James J. Glackin. Letters: Peaceful England? Bulletin of the Atomic Scientists, 31(7):3, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mel75].

Glackin:1976:DDU

[Gla76a] James J. Glackin. The dangerous drift in uranium enrichment. Bulletin of the Atomic Scientists, 32(2):22–29, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Roc76].

Glackin:1976:LCD

[Gla76b] James J. Glackin. Letter: Courting destruction. Bulletin of the Atomic Scientists, 32(3):3–4, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75b].

Glackin:1976:LTC

[Gla76c] James J. Glackin. Letter: Twentieth-century feudalism. Bulletin of the Atomic Scientists, 32(10):2, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sal76].

Glackin:1976:LGA

[Gla76d] James L. Glackin. Letter: Guilt by association. Bulletin of the Atomic Scientists, 32(5):4, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch75e].

Glamann:1976:NM

[Gla76e] Daryl Glamann, Sr. Nuclear moratorium. Bulletin of the Atomic Scientists, 32(10):4, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glass:1978:LLC

[Gla78] Jonathan Clyde Glass. Letter: Livermore conflict. Bulletin of the Atomic Scientists, 34(6):4–5, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glaser:1985:SWB

[Gla85] Charles L. Glaser. Star Wars bad even if it works. Bulletin of the Atomic Scientists, 41(3):13–16, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Abl85].

Glaser:2002:WUB

[Gla02] Alexander Glaser. Weapons uranium: Bavaria bucks ban. Bulletin of the Atomic Scientists, 58(2):20–22, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/2/20.

Glaser:2012:BFL

[Gla12] Alexander Glaser. From Brokdorf to Fukushima: The long journey to nuclear phase-out. Bulletin of the Atomic Scientists, 68(6): 10–21, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/68/6/10.

Glennan:1951:NHI

[Gle51] T. Keith Glennan. New horizons for industry in the isotopes field. Bulletin of the Atomic Scientists, 7(6):174–176, 191, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glennan:1952:LEB

[Gle52] T. Keith Glennan. Letter to the Editor: Where Credit is Due. Bulletin of the Atomic Scientists, 8(2):65, February 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gleditsch:1983:EFN

[Gle83] Nils Petter Gleditsch. From Europe: The freeze in Norway. Bulletin of the Atomic Scientists, 39(9):32–34, November 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gleick:1991:ESC

[Gle91] Peter H. Gleick. Environment and security: The clear connections. Bulletin of the Atomic Scientists, 47(3):16–21, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bye91].

Gleick:1994:BRW

[Gle94] Peter Gleick. Book review: Whole-earth security: Ultimate Security: How Environmental Concerns Affect Global Political Stability, by Norman Myers. Bulletin of the Atomic Scientists, 50 (2):55–56, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gleick:2021:WRN

[Gle21] Peter H. Gleick. Water recommendation for the new administration. Bulletin of the Atomic Scientists, 77(1):24–27, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1860326.

Gleick:2024:PWE

[Gle24] Peter H. Gleick. Peak water in an era of climate change. Bulletin of the Atomic Scientists, 80(4):213–217, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2364505.

Gard:1979:CHN

[GLH^{+}79] Robert G. Gard, Jr., Gene R. La Rocque, John T. Hayward, Steve Hoffman, Michael Sullivan, Herbert Scoville, A. S. Collins, Jr., Richard Barnet, Rudolph Raisin, and Chris Anfinsen. Comments: [on how a nuclear war would be fought]. Bulletin of the Atomic Scientists, 35(5):30–33, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gregory:1960:RCR

[GLMS60] Bernard Gregory, Antoine Lacassagne, Herbert Marcovich, and Édouard Selzer. Readers’ comment: Regarding possible scientific value of further nuclear explosions. Bulletin of the Atomic Scientists, 16(3):110, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldemberg:2007:WCW

[GLSS07] Jose Goldemberg, Amory B. Lovins, Stephen Schneider, and M. S. Swaminathan. What can we do? Bulletin of the Atomic Scientists, 63(1):47–48, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/47.

Glushchenko:1950:SSU

[Glu50] I. E. Glushchenko. Science in the Soviet Union: a Soviet professor’s letter. Bulletin of the Atomic Scientists, 6(6):178–179, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1954:RSP

[GM54] Gordon Gray and Thomas A. Morgan. Report of the Special Personnel Security Board: Majority report, May 27, 1954. Bulletin of the Atomic Scientists, 10(6):243–249, 254, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Galpern:1987:LDR

[GM87] Daniel M. Galpern and Gerald E. Marsh. Letters: Don’t rely on “Navy’s MX”. Bulletin of the Atomic Scientists, 43(9):53– 54, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar87a, Mar97a].

Gould:1990:LWB

[GM90] Jay Gould and Paul Meier. Letters: Why is baby-boomer death rate increasing? Bulletin of the Atomic Scientists, 46(10):50– 51, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mei90].

Gupta:1993:SH

[GM93] Vipin Gupta and Philip McNab. Sleuthing from home. Bulletin of the Atomic Scientists, 49(10):44–47, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Glaser:2009:FMS

[GM09a] Alexander Glaser and Zia Mian. Fissile material stocks and production, 2008. Bulletin of the Atomic Scientists, 65(1):35–47, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/65/1/35.

Goodby:2009:RRN

[GM09b] James Goodby and Fred McGoldrick. Reducing the risks of nuclear power’s global spread. Bulletin of the Atomic Scientists, 65(3):40–47, May 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/65/3/40.

Gilinsky:2010:RNA

[GM10] Victor Gilinsky and Roger J. Mattson. Revisiting the NUMEC Affair. Bulletin of the Atomic Scientists, 66(2):61–75, March 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/2/61.

Gronlund:2014:AAN

[GM14] Lisbeth Gronlund and Eryn MacDonald. America, awash in nuclear weapons materials. Bulletin of the Atomic Scientists, 70(2): 53–67, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gregg:2024:GCN

[GMV^{+}24] Daniel J. Gregg, John S. McCloy, John D. Vienna, Allison M. Macfarlane, William J. Weber, and Gregory R. Lumpkin. Glass and ceramic nuclear waste forms: The scientific battle. Bulletin of the Atomic Scientists, 81(1):3–16, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2441044.

Ghoshroy:2010:SAC

[GN10] Subrata Ghoshroy and Götz Neuneck. South Asia at a crossroads: conflict or cooperation in the age of nuclear weapons, missile defense, and space rivalries, volume 197 of Demokratie, Sicherheit, Frieden; Democracy, Security, Peace; v. 197. Nomos, Baden-Baden, Germany, 2010. ISBN 3-8329-5398-1. 330 pp. LCCN JZ6009.S64 S68 2010.

Gnepp:1977:LBP

[Gne77] Armen Gnepp. Letter: Bold proposal. Bulletin of the Atomic Scientists, 33(10):9, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fin77a].

Garrison:1954:MJR

[GNMS54] Lloyd K. Garrison, General K. D. Nichols, Thomas E. Murray, and Henry DeWolf Smyth. “In the matter of J. Robert Oppenheimer”: Mr. Garrison to General Nichols; findings and recommendations of General Nichols. Bulletin of the Atomic Scientists, 10(7):270–274, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goens:1979:CNP

[Goe79] J. Goens. Commentary: Nuclear power/nuclear weapons. Bulletin of the Atomic Scientists, 35(10):57, December 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goen:1980:CDM

[Goe80] Richard L. Goen. Commentary: A dangerous myth. Bulletin of the Atomic Scientists, 36(1):63, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79a].

Gofman:1971:NPE

[Gof71] John W. Gofman. Nuclear power and ecocide: An adversary view of new technology. Bulletin of the Atomic Scientists, 27(7):28– 32, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gofman:1993:BDD

[Gof93a] John W. Gofman. Beware the data diddlers. Bulletin of the Atomic Scientists, 49(4):40–44, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gofman:1993:RR

[Gof93b] John W. Gofman. The rules of research. Bulletin of the Atomic Scientists, 49(4):42, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gofman:1993:VRR

[Gof93c] John W. Gofman. Violating the rules of research. Bulletin of the Atomic Scientists, 49(4):43, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldhamer:1948:QMR

[Gol48] Herbert Goldhamer. A question of moral responsibility. Bulletin of the Atomic Scientists, 4(3):73–74, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bri48a].

Goldschmidt:1949:RP

[Gol49] Richard B. Goldschmidt. Research and politics. Bulletin of the Atomic Scientists, 5(5):150–155, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberger:1955:BBC

[Gol55] Marvin L. Goldberger. Books: Science and the Common Understanding, by J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 11(2):64, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldbert:1961:S

[Gol61a] Leo Goldbert. The sun. Bulletin of the Atomic Scientists, 17(5– 6):210–213, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldsmith:1961:UAC

[Gol61b] John R. Goldsmith. Urban air conservation. Bulletin of the Atomic Scientists, 17(9):376–379, 386, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1961:PKU

[Gol61c] Walter Goldstein. The price of keeping the U.N. Bulletin of the Atomic Scientists, 17(8):315–321, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldschmidt:1962:FAE

[Gol62a] Bertrand Goldschmidt. The French atomic energy program. Bulletin of the Atomic Scientists, 18(8):46–48, October 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldschmidt:1962:NRF

[Gol62b] Bertrand Goldschmidt. News & reports: The French atomic energy program. Bulletin of the Atomic Scientists, 18(7):39–42, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1962:UD

[Gol62c] Walter Goldstein. The U.N. and its detractors. Bulletin of the Atomic Scientists, 18(4):12–17, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Golovin:1966:TCP

[Gol66] Nicholas E. Golovin. The N th country’s problem in space exploration. Bulletin of the Atomic Scientists, 22(10):13–18, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1967:BBA

[Gol67a] Walter Goldstein. Books: The Arrogance of Power, by Senator J. William Fulbright. Bulletin of the Atomic Scientists, 23(8): 30–32, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1967:DNP

[Gol67b] Walter Goldstein. A dynamic new policy toward the new Europe. Bulletin of the Atomic Scientists, 23(4):17–22, April 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldhaber:1969:TME

[Gol69a] Michael H. Goldhaber. Technological man: Exploding the myths? Bulletin of the Atomic Scientists, 25(9):41–42, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1969:BBM

[Gol69b] Walter Goldstein. Books: No More Vietnams? The War and the Future of American Foreign Policy, by Richard M. Pfeffer; Intervention and Revolution: The United States in the Third World, by Richard J. Barnet. Bulletin of the Atomic Scientists, 25(10):39– 40, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldwasser:1969:SMB

[Gol69c] Edwin L. Goldwasser. Science and man: Breaking new ground at Batavia. Bulletin of the Atomic Scientists, 25(8):7–10, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1970:CTG

[Gol70a] Jozef Goldblat. Comment: Are tear gas and herbicides permitted weapons? Bulletin of the Atomic Scientists, 26(4):13–16, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldsmith:1970:CA

[Gol70b] Maurice Goldsmith. Crisis in Aspen. Bulletin of the Atomic Scientists, 26(9):28–30, 45, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1970:LVW

[Gol70c] Walter Goldstein. The lessons of the Vietnam War. Bulletin of the Atomic Scientists, 26(2):41–45, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1972:BD

[Gol72a] Jozef Goldblat. Biological disarmament. Bulletin of the Atomic Scientists, 28(4):6–10, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldie:1972:CPT

[Gol72b] L. F. E. Goldie. Communications: Pollution taxes. Bulletin of the Atomic Scientists, 28(2):3–4, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ram70].

Goldsmith:1973:LLM

[Gol73a] Maurice Goldsmith. Letter from London: Meadows unlimited or caveat computer. Bulletin of the Atomic Scientists, 29(5):16–18, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldsmith:1973:LLT

[Gol73b] Maurice Goldsmith. Letter from London: The thalidomide affair. Bulletin of the Atomic Scientists, 29(3):40–41, March 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gold:1974:SIW

[Gol74] T. Gold. Skylab: Is it worth the risk and the expense? Bulletin of the Atomic Scientists, 30(2):4–8, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1975:LSR

[Gol75] Lazer Goldberg. Letters: The scientist’s role. Bulletin of the Atomic Scientists, 31(5):2, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel75].

Goldstein:1977:LPS

[Gol77] Martin E. Goldstein. Letter: Pentagon seeks help. Bulletin of the Atomic Scientists, 33(2):2, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Golden:1978:BBS

[Gol78a] William T. Golden. Books: Sputnik, Scientists, and Eisenhower: A Memoir of the First Special Assistant to the President for Science and Technology, by James R. Killian, Jr. Bulletin of the Atomic Scientists, 34(4):60–62, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1978:PFU

[Gol78b] Walter Goldstein. The political failure of U.S. energy policy. Bulletin of the Atomic Scientists, 34(9):17–19, November 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Golsong:1978:ECH

[Gol78c] Heribert Golsong. European Convention on Human Rights. Bulletin of the Atomic Scientists, 34(8):7–8, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1979:CCE

[Gol79] Walter Goldstein. Commentary: Carter’s energy program stalls in Congress. Bulletin of the Atomic Scientists, 35(9):7–10, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldemberg:1982:SHR

[Gol82a] José Goldemberg. Scientists and human rights in Latin America. Bulletin of the Atomic Scientists, 38(1):35–38, January 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldhaber:1982:CC

[Gol82b] Maurice Goldhaber. With Chadwick at the Cavendish. Bulletin of the Atomic Scientists, 38(10):12–13, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1982:SUE

[Gol82c] Marshall I. Goldman. The Soviet Union: The economy and the consumer. Bulletin of the Atomic Scientists, 38(4):24–27, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1983:CIW

[Gol83a] Jozef Goldblat. The convention on ‘inhumane’ weapons. Bulletin of the Atomic Scientists, 39(1):24–25, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1983:TF

[Gol83b] Jozef Goldblat. Tlatelolco and the Falklands. Bulletin of the Atomic Scientists, 39(5):49, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldsmith:1983:BRB

[Gol83c] George J. Goldsmith. Book review: Nuclear Hostages, by Bernard J. O’Keefe. Bulletin of the Atomic Scientists, 39(9):40–41, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1984:CTV

[Gol84a] Jozef Goldblat. Charges of treaty violations. Bulletin of the Atomic Scientists, 40(5):33–36, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1984:CTS

[Gol84b] Louis Goldman. Commentary: Is there a Soviet bomb in Wichita? Bulletin of the Atomic Scientists, 40(1):54–56, January 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1985:CWV

[Gol85a] Jozef Goldblat. Chemical weapons verification. Bulletin of the Atomic Scientists, 41(5):19, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1985:WEG

[Gol85b] Marshall I. Goldman. What to expect from Gorbachev. Bulletin of the Atomic Scientists, 41(5):8–9, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1986:WNS

[Gol86a] Jozef Goldblat. Will the NPT survive? Bulletin of the Atomic Scientists, 42(1):35–38, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gollon:1986:SFC

[Gol86b] Peter J. Gollon. SDI funds costly for scientists. Bulletin of the Atomic Scientists, 42(1):24–26, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sam86].

Goldberg:1988:BRBa

[Gol88a] Andrew Goldberg. Book review: Wargames: The Secret World of the Creators, Players and Policymakers Rehearsing World War III Today, by Thomas B. Allen. Bulletin of the Atomic Scientists, 44(5):48–50, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1988:BRBb

[Gol88b] Stanley Goldberg. Book review: The General and the Bomb: A Biography of General Leslie R. Groves, Director of the Manhattan Project, by William Lawren. Bulletin of the Atomic Scientists, 44 (10):44–45, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1988:LCA

[Gol88c] Jozef Goldblat. Letter: Correct antecedent. Bulletin of the Atomic Scientists, 44(6):54, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol88d].

Goldblat:1988:LSP

[Gol88d] Jozef Goldblat. Letter: Submarine proliferation. Bulletin of the Atomic Scientists, 44(4):53, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim87b] and correction [Gol88c].

Goldblat:1988:LVU

[Gol88e] Jozef Goldblat. Letter: Verification not U.N. task. Bulletin of the Atomic Scientists, 44(10):48–49, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AGC^{+}88] and comment [Hei89].

Goldblat:1988:WIW

[Gol88f] Jozef Goldblat. What it would take to ban testing. Bulletin of the Atomic Scientists, 44(8):25–27, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstein:1988:BRB

[Gol88g] Gary R. Goldstein. Book review: Star Wars: The Economic Fallout, by the Council on Economic Priorities. Bulletin of the Atomic Scientists, 44(4):47–48, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldstene:1989:LSV

[Gol89] Paul N. Goldstene. Letter: Save the voice. Bulletin of the Atomic Scientists, 45(3):44, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar89a].

Goldblat:1990:LRD

[Gol90a] Jozef Goldblat. Letter: Reactor deals flout NPT. Bulletin of the Atomic Scientists, 46(5):43–44, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sco90].

Goldemberg:1990:POK

[Gol90b] Jose Goldemberg. Perspective: One kilowatt per capita. Bulletin of the Atomic Scientists, 46(1):13–14, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1990:RAU

[Gol90c] Joe Goldman. Report: Argentina: U.S. endorses Menem’s nuclear plans. Bulletin of the Atomic Scientists, 46(6):9–10, July/ August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1990:RNPa

[Gol90d] Joe Goldman. Report: Nuclear power: Bad news in Buenos Aires. Bulletin of the Atomic Scientists, 46(5):8–9, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1990:RNPb

[Gol90e] Joe Goldman. Report: Nuclear power: Chile takes the open road. Bulletin of the Atomic Scientists, 46(7):7–8, September 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gold:1991:MRP

[Gol91a] David Gold. Military R&D a poor scapegoat for flagging economy. Bulletin of the Atomic Scientists, 47(1):38–43, January 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1991:BAC

[Gol91b] Joe Goldman. Bulletin: Armed capitalism. Bulletin of the Atomic Scientists, 47(4):5–6, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldman:1991:RNA

[Gol91c] Joe Goldman. Report: Nonproliferation: Argentina, Brazil open to inspections. Bulletin of the Atomic Scientists, 47(4):8–10, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gold:1992:BRG

[Gol92a] David Gold. Book review: Guns and butter: The Rise of the Gunbelt: The Remapping of Industrial America, by Ann Markusen, Peter Hall, Scott Campbell, and Sabine Deitrick. Bulletin of the Atomic Scientists, 48(3):44–45, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:GSO

[Gol92b] Stanley Goldberg. Groves, Szilard, and Oppenheimer. Bulletin of the Atomic Scientists, 48(10):37, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:GTR

[Gol92c] Stanley Goldberg. Groves takes the reins. Bulletin of the Atomic Scientists, 48(10):32–39, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gold:1993:BRR

[Gol93a] David Gold. Book review: Reducing R&D: Dismantling the Cold War Economy, by Ann Markusen and Joel Yudkin. Bulletin of the Atomic Scientists, 49(4):45–46, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldanskii:1993:LSE

[Gol93b] Vitalii I. Goldanskii. Letter: Scientists excepted. Bulletin of the Atomic Scientists, 49(8):54, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol93c].

Goldanskii:1993:RRB

[Gol93c] Vitalii I. Goldanskii. Russia’s “red–brown” hawks. Bulletin of the Atomic Scientists, 49(5):24–27, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See author’s objection to article title [Gol93b].

Goldfarb:1993:BRF

[Gol93d] Jeffrey C. Goldfarb. Book review: Free to falter: Reinventing Politics: Eastern Europe from Stalin to Havel, by Vladimir Tismaneanu. Bulletin of the Atomic Scientists, 49(2):44, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldring:1994:RNS

[Gol94] Natalie J. Goldring. Report: NATO: Skittish on counterproliferation. Bulletin of the Atomic Scientists, 50(2):12–13, March/ April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1995:BRC

[Gol95a] Stanley Goldberg. Book review: The Conant conundrum: James B. Conant: Harvard to Hiroshima and the Making of the Nuclear Age, by James G. Hershberg. Bulletin of the Atomic Scientists, 51 (2):63–65, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1995:GSC

[Gol95b] Stanley Goldberg. Groves and the scientists: Compartmentalization and the building of the bomb. Physics Today, 48(8):38–43, August 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v48/i8/p38/s1.

Goldberg:1995:SSL

[Gol95c] Stanley Goldberg. Smithsonian suffers Legionnaires’ disease. Bulletin of the Atomic Scientists, 51(3):28–33, May/June 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HWG95, Cam95b, Col95].

Goldring:1995:BRP

[Gol95d] Natalie J. Goldring. Book review: Patterns of death: The Global Spread of Arms: Political Economy of International Security, by Frederic S. Pearson. Bulletin of the Atomic Scientists, 51(5):57– 58, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1996:BRA

[Gol96a] Stanley Goldberg. Book review: Another death greatly exaggerated: The End of Science: Facing the Limits of Knowledge in the Twilight of the Scientific Age, by John Horgan. Bulletin of the Atomic Scientists, 52(6):60–62, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1996:BRF

[Gol96b] Stanley Goldberg. Book review: With friends like these …: Lise Meitner: A Life in Physics, by Ruth Lewin Sime. Bulletin of the Atomic Scientists, 52(4):55–57, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldblat:1997:LMP

[Gol97] Jozef Goldblat. Letter: Mandatory penalties. Bulletin of the Atomic Scientists, 53(5):60–61, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1998:WDT

[Gol98] Stanley Goldberg. Perspective: What did Truman know, and when did he know it? Bulletin of the Atomic Scientists, 54(3):18– 19, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldring:1999:NGG

[Gol99] Natalie J. Goldring. The NRA goes global. Bulletin of the Atomic Scientists, 55(1):61–65, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/61.

Goldring:2007:CW

[Gol07] Natalie J. Goldring. Conventional wisdom. Bulletin of the Atomic Scientists, 63(2):58–60, March/April 2007. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/2/58.

Goldstein:2009:HLW

[Gol09] Jeff Goldstein. How light water reactors figure into negotiations with North Korea. Bulletin of the Atomic Scientists, 65(4):64–71, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/4/64.

Goldstein:2011:HBB

[Gol11] Jon Goldstein. How to build a better sepulcher: Lessons from new Mexico’s waste isolation pilot plant. Bulletin of the Atomic Scientists, 67(5):77–88, September/October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/5/77.

Goldman:2021:DPB

[Gol21] Eric Goldman. Dear President Biden: You should save, not revoke, Section 230. Bulletin of the Atomic Scientists, 77(1):36– 37, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1859863.

Goldstein:2023:TT

[Gol23] Lyle Goldstein. The trouble with Taiwan. Bulletin of the Atomic Scientists, 79(2):103–107, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2178707.

Goldston:2024:OFL

[Gol24] Robert J. Goldston. An overview of the fusion landscape. Bulletin of the Atomic Scientists, 80(6):365–370, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2417611.

Gomer:1961:STA

[Gom61a] Robert Gomer. Some thoughts on arms control. Bulletin of the Atomic Scientists, 17(4):133–137, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1961:STC

[Gom61b] Robert Gomer. Some thoughts on China. Bulletin of the Atomic Scientists, 17(5–6):241–243, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1962:LEE

[Gom62a] Robert Gomer. Letter to the Editor: Erratum. Bulletin of the Atomic Scientists, 18(7):38, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom62c].

Gomer:1962:LEV

[Gom62b] Robert Gomer. Letter to the Editor: Vox populi. Bulletin of the Atomic Scientists, 18(9):30–31, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom62c].

Gomer:1962:VP

[Gom62c] Robert Gomer. Vox populi. Bulletin of the Atomic Scientists, 18(6):32–33, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gom62b] and erratum [Gom62a].

Gomer:1965:ADS

[Gom65] Robert Gomer. The ABM debate: a Soviet view. Bulletin of the Atomic Scientists, 21(2):25–26, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1967:AD

[Gom67] Robert Gomer. The ABM decision. Bulletin of the Atomic Scientists, 23(9):29, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1968:TP

[Gom68] Robert Gomer. The tyranny of progress. Bulletin of the Atomic Scientists, 24(2):4–8, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1970:SPA

[Gom70] Robert Gomer. Science and public affairs communications: Letter from Paris: The quality of life. Bulletin of the Atomic Scientists, 26(2):55–56, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1971:BBL

[Gom71] Robert Gomer. Books: The Lost Crusade — America in Vietnam, by Chester Cooper. Bulletin of the Atomic Scientists, 27 (5):45–46, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1974:IED

[Gom74] Robert Gomer. India explodes a ‘device’. Bulletin of the Atomic Scientists, 30(6):3, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pya74].

Gomer:1975:NWPa

[Gom75a] Robert Gomer. Notes on the world: Part 2: Europe and Asia through American eyes: From the Greek Islands to Katmandu. Bulletin of the Atomic Scientists, 31(4):35–41, April 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Wag75, Ris75, Zel75].

Gomer:1975:NWPb

[Gom75b] Robert Gomer. Notes on the world, Part 3: The importance of point of view: Hong Kong, Japan and home. Bulletin of the Atomic Scientists, 31(5):34–39, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zel75].

Gomer:1975:SNW

[Gom75c] Robert Gomer. Some notes on the world: Trieste and the Alps; the Germans and the Austrians. Bulletin of the Atomic Scientists, 31(3):34–40, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zel75].

Gomberg:1978:LPF

[Gom78] Henry J. Gomberg. Letter: Peacetime fusion. Bulletin of the Atomic Scientists, 34(6):3, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol78].

Gomer:1980:CET

[Gom80] Robert Gomer. Commentary: Election thoughts. Bulletin of the Atomic Scientists, 36(8):11–12, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1981:CCE

[Gom81a] Robert Gomer. [Comment on ‘A critique of the END campaign’]. Bulletin of the Atomic Scientists, 37(10):38, December 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1981:CB

[Gom81b] Robert Gomer. Commentary: Baghdad. Bulletin of the Atomic Scientists, 37(7):7–8, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1981:VH

[Gom81c] Robert Gomer. Vanishing horizons. Bulletin of the Atomic Scientists, 37(2):10–13, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1982:LL

[Gom82a] Robert Gomer. At long last. Bulletin of the Atomic Scientists, 38(6):3, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1982:FYA

[Gom82b] Robert Gomer. Forty years ago. Bulletin of the Atomic Scientists, 38(10):10–11, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1983:PUS

[Gom83] Robert Gomer. Perspective: The United States and Central America. Bulletin of the Atomic Scientists, 39(7):3–4, August/ September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1984:BRB

[Gom84a] Robert Gomer. Book review: Weapons and Hope, by Freeman Dyson. Bulletin of the Atomic Scientists, 40(8):36–40, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1984:CW

[Gom84b] Robert Gomer. Commentary: Where to? Bulletin of the Atomic Scientists, 40(2):46–47, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [PG84].

Gomez:1987:LST

[Gom87] Artie Gomez. Letter: “Star Trek” is not Star Wars. Bulletin of the Atomic Scientists, 43(3):52–53, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi87c] and comment [Smi87b].

Gompert:2000:SF

[Gom00] David C. Gompert. Sharpen the fear. Bulletin of the Atomic Scientists, 56(1):22–23, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/22.

Goodwin:1955:AB

[Goo55] Irwin Goodwin. The AAAS at Berkeley. Bulletin of the Atomic Scientists, 11(3):82–84, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goodall:1977:LNB

[Goo77] M. C. Goodall. Letter: No bargain. Bulletin of the Atomic Scientists, 33(2):6, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei76a].

Goodman:1984:BRB

[Goo84] Joseph C. Goodman. Book review: The Media Monopoly, by Ben H. Bagdikian. Bulletin of the Atomic Scientists, 40(1):38– 39, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Good:1985:AUP

[Goo85a] Paul Good. Athletes unite for peace. Bulletin of the Atomic Scientists, 41(10):43–44, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goodby:1985:UVS

[Goo85b] James E. Goodby. U.S. view on Stockholm conference. Bulletin of the Atomic Scientists, 41(3):35–36, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goose:1986:BRB

[Goo86] Stephen D. Goose. Book reviews: World Military Expenditures and Arms Transfer 1985; World Military and Social Expenditures 1985, by Ruth Leger Sivard. Bulletin of the Atomic Scientists, 42(2):43–46, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goodman:2000:MCP

[Goo00] Melvin A. Goodman. The milk churn papers. Bulletin of the Atomic Scientists, 56(3):64–65, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/3/64.

Goodrich:2003:TDA

[Goo03] Peter Spang Goodrich. Targeting: Dead aim still deadly. Bulletin of the Atomic Scientists, 59(4):18–19, July/August 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/4/18.

Goodchild:2004:ETRa

[Goo04a] Peter Goodchild. Edward Teller: the real Dr Strangelove. Weidenfeld and Nicolson, London, UK, 2004. ISBN 0-297-60734-0. 467 pp. LCCN QC16.T37 G66 2004b.

Goodchild:2004:ETRb

[Goo04b] Peter Goodchild. Edward Teller, the real Dr. Strangelove. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01669-6. xxv + 469 + 16 pp. LCCN QC16.T37 G66 2004. URL

http://alsos.wlu.edu/information.aspx?id=2237.

Goodchild:2005:SHP

[Goo05] Peter Goodchild. The sum of his parts. Bulletin of the Atomic Scientists, 61(6):63–66, November/December 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/6/63;
 http://bos.sagepub.com/content/61/6/63.abstract. Re-

view of [BS05, McM05].

Gorrie:1951:FDP

[Gor51a] Jack Gorrie. Federal dispersion policy stresses local initiative. Bulletin of the Atomic Scientists, 7(9):270–271, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gorrie:1951:ODR

[Gor51b] Jack Gorrie. Official dispersal recommendations of the National Security Resources Board: Foreword. Bulletin of the Atomic Scientists, 7(9):264, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1959:NMS

[Gor59] Bernard K. Gordon. NATO’s missing shield. Bulletin of the Atomic Scientists, 15(6):229–233, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1960:TPH

[Gor60] Bernard K. Gordon. “The Top of Policy Hill”. Bulletin of the Atomic Scientists, 16(7):289–291, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gorove:1967:COS

[Gor67a] Stephen Gorove. Comments: The outer space treaty. Bulletin of the Atomic Scientists, 23(10):44–45, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gorove:1967:ICE

[Gor67b] Stephen Gorove. Inspection and control in Euratom. Bulletin of the Atomic Scientists, 23(3):41–46, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gorr:1971:BBS

[Gor71] Louis F. Gorr. Books: The Social Responsibility of the Scientist, by Martin Brown. Bulletin of the Atomic Scientists, 27(6):36–37, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gorr:1973:BBT

[Gor73] Louis F. Gorr. Books: Think Tanks, by Paul Dickson. Bulletin of the Atomic Scientists, 29(5):46–47, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goran:1976:SAH

[Gor76] Morris Goran. The scientist as anti-hero. Bulletin of the Atomic Scientists, 32(3):23, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1980:CRP

[Gor80] Bruce Gordon. Commentary: Registering a protest. Bulletin of the Atomic Scientists, 36(8):64, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1981:EDC

[Gor81] Lincoln Gordon. Energy development: crisis and transition. Bulletin of the Atomic Scientists, 37(4):24–29, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gore:1984:PBA

[Gor84] Albert Gore, Jr. Perspective: A bipartisan approach to arms control. Bulletin of the Atomic Scientists, 40(8):7–8, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hel85].

Gordon:1985:LSS

[Gor85] Arnold Gordon. Letter: Sakharov and Soviet emigration. Bulletin of the Atomic Scientists, 41(2):50, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu84b].

Gorbachev:1992:RT

[Gor92] Mikhail Gorbachev. The river of time. Bulletin of the Atomic Scientists, 48(6):22–27, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1995:CFA

[Gor95a] Danielle Gordon. Cut foreign arms sales. Bulletin of the Atomic Scientists, 51(5):47–48, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1995:MM

[Gor95b] Danielle Gordon. The MILSTAR millstone. Bulletin of the Atomic Scientists, 51(5):50, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1995:MMF

[Gor95c] Danielle Gordon. More missiles, fewer targets. Bulletin of the Atomic Scientists, 51(5):47, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1995:PW

[Gor95d] Danielle Gordon. Prosperity for whom? Bulletin of the Atomic Scientists, 51(5):43, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1995:ROS

[Gor95e] Danielle Gordon. Readiness and other strawmen. Bulletin of the Atomic Scientists, 51(5):40–41, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1995:UO

[Gor95f] Danielle Gordon. Underfunding? Or overprogramming? Bulletin of the Atomic Scientists, 51(5):38, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1996:BGR

[Gor96a] Danielle Gordon. Bulletin: Getting a real bang for the buck. Bulletin of the Atomic Scientists, 52(4):6, July/August 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1996:BNP

[Gor96b] Danielle Gordon. Bulletin: No peace for South Africa’s wand’ring warriors. Bulletin of the Atomic Scientists, 52(3):5–6, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1996:BWC

[Gor96c] Danielle Gordon. Bulletin: Where in the (cyber) world is Carlos Salinas. Bulletin of the Atomic Scientists, 52(5):6, September/ October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1996:CSS

[Gor96d] Danielle Gordon. Coming soon to a sky near you. Bulletin of the Atomic Scientists, 52(5):50, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1996:EC

[Gor96e] Danielle Gordon. Experimentation continues. Bulletin of the Atomic Scientists, 52(1):36, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gordon:1996:VNH

[Gor96f] Danielle Gordon. The verdict: No harm, no foul. Bulletin of the Atomic Scientists, 52(1):32–40, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [JG96].

Gormley:2006:CC

[Gor06] Dennis M. Gormley. Cruise control. Bulletin of the Atomic Scientists, 62(2):26–33, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/2/26.

Gorbachev:2011:CYL

[Gor11] Mikhail Gorbachev. Chernobyl 25 years later: Many lessons learned. Bulletin of the Atomic Scientists, 67(2):77–80, March/ April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/67/2/77.

Gorbachev:2020:CYL

[Gor20] Mikhail Gorbachev. 2011: Chernobyl 25 years later: Many lessons learned. Bulletin of the Atomic Scientists, 76(6):428– 430, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00963402.2020.1847523.

Gotlieb:1950:CAI

[Got50] Albert Gotlieb. The Council on Atomic Implications. Bulletin of the Atomic Scientists, 6(2):61, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gotlieb:1951:CAI

[Got51] Albert Gotlieb. The Council on Atomic Implications. Bulletin of the Atomic Scientists, 7(1):27, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottlieb:1962:CP

[Got62] Sanford Gottlieb. Campaigning for peace. Bulletin of the Atomic Scientists, 18(9):38–39, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottlieb:1964:BBS

[Got64] Sanford Gottlieb. Books: Professional Staffs of Congress, by Kenneth Kofmehl. Bulletin of the Atomic Scientists, 20(4):36–37, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottlieb:1965:CR

[Got65] Sanford Gottlieb. Conversing with Russians. Bulletin of the Atomic Scientists, 21(1):35–36, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gotaas:1969:OPA

[Got69] Harold B. Gotaas. Outwitting the patient assassin: The human use of lake pollution. Bulletin of the Atomic Scientists, 25(5): 8–10, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottstein:1977:NET

[Got77] Klaus Gottstein. Nuclear energy and the Third World. Bulletin of the Atomic Scientists, 33(6):44–48, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottstein:1982:EM

[Got82] Klaus Gottstein. Escaping the minefield. Bulletin of the Atomic Scientists, 38(8):55–56, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottstein:1983:CGE

[Got83] Klaus Gottstein. Commentary: The Göttingen Eighteen. Bulletin of the Atomic Scientists, 39(8):62–63, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottlieb:1987:BRB

[Got87] Sanford Gottlieb. Book review: Blundering into Disaster: Surviving the First Century of the Nuclear Age, by Robert S. Mc-Namara. Bulletin of the Atomic Scientists, 43(2):56–57, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gottemoeller:2020:SDE

[Got20] Rose Gottemoeller. Science diplomacy: The essential interdisciplinary approach. Bulletin of the Atomic Scientists, 76(6): 321–324, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1846419.

Goudsmit:1946:HGL

[Gou46] Samuel A. Goudsmit. How Germany lost the race. Bulletin of the Atomic Scientists, 1(7):4–5, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1947:GSA

[Gou47a] Samuel A. Goudsmit. German scientists in Army employment: I. The case analyzed. Bulletin of the Atomic Scientists, 3(2):64, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1947:HGN

[Gou47b] Samuel A. Goudsmit. Heisenberg on the German Nuclear Power Project. Bulletin of the Atomic Scientists, 3(11):343, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Hei47].

Goudsmit:1948:OTG

[Gou48] Samuel A. Goudsmit. Our task in Germany. Bulletin of the Atomic Scientists, 4(4):106, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1955:TSO

[Gou55] S. A. Goudsmit. The task of the security officer. Bulletin of the Atomic Scientists, 11(4):145–147, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Goudsmit:1959:LEF

[Gou59] S. A. Goudsmit. Letter to the Editor: Fallout fallacies. Bulletin of the Atomic Scientists, 15(7):321, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1961:LEB

[Gou61] S. A. Goudsmit. Letter to the Editor: Big risks and small. Bulletin of the Atomic Scientists, 17(9):393, November 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1963:LEC

[Gou63] S. A. Goudsmit. Letter to the Editor: “Credo of an angry old man”. Bulletin of the Atomic Scientists, 19(5):33–34, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gould:1970:EAM

[Gou70] Laurence M. Gould. Emergence of Antarctica: The mythical land. Bulletin of the Atomic Scientists, 26(10):5–10, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Goudsmit:1974:CS

[Gou74] S. A. Goudsmit. Communications: I am scared. Bulletin of the Atomic Scientists, 30(2):2–3, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author writes about the failure of the German atomic bomb project in 1944, and American military attitudes about use of atomic weapons. See comment [Sau74].

Goulet:1975:LCV

[Gou75a] Denis Goulet. Letter: Correction [on views of Franklin Tugwell]. Bulletin of the Atomic Scientists, 31(10):4, December 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goulet:1975:PTT

[Gou75b] Denis Goulet. The paradox of technology transfer: The Third World’s ladder to development can also be the avenue to colonialism. Bulletin of the Atomic Scientists, 31(6):39–46, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction on view of Franklin Tugwell [Gou75a].

Goudsmit:1977:ID

[Gou77] Samuel A. Goudsmit. An irrelevant debate. Bulletin of the Atomic Scientists, 33(3):67, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goudsmit:1978:LNG

[Gou78a] S. A. Goudsmit. Letter: Numbers game. Bulletin of the Atomic Scientists, 34(1):6–7, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tho77].

Goudsmit:1978:BSU

[Gou78b] Samuel Goudsmit. Books: Scientists under Hitler: Politics and the Physics Community in the Third Reich, by Alan D. Beyerchen. Bulletin of the Atomic Scientists, 34(7):47–49, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goure:1978:AI

[Gou78c] Leon Gouré. Another interpretation. Bulletin of the Atomic Scientists, 34(4):48–51, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Kap78a].

Gould:1995:LBB

[Gou95] Jay M. Gould. Letter: Of bombs and biology. Bulletin of the Atomic Scientists, 51(1):68, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol94b].

Goudarzi:2023:PCH

[Gou23] Sara Goudarzi. Popping the chatbot hype balloon. Bulletin of the Atomic Scientists, 79(5):293–298, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2245244.

Goudarzi:2024:PDV

[Gou24] Sara Goudarzi. To protect democratic values, journalism must save itself. Bulletin of the Atomic Scientists, 80(5):314–320, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2388468.

Govan:1993:BRD

[Gov93] Emilia L. Govan. Book review: Danger: defense droppings: The Threat at Home: The Toxic Legacy of the U.S. Military, by Seth Shulman. Bulletin of the Atomic Scientists, 49(1):49–50, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gowing:1979:RAE

[Gow79] Margaret Gowing. Reflections on atomic energy history. Bulletin of the Atomic Scientists, 35(3):51–54, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Based on the 1978 Rede Lecture at Cambridge University.

Garwin:1984:SWH

[GP84] Richard L. Garwin and John Pike. Space weapons: History and current debate. Bulletin of the Atomic Scientists, 40(5):[various between 32–33], May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [OGP84].

Goldberg:1992:ATG

[GP92a] Stanley Goldberg and Thomas Powers. Alsos: Tracking the German bomb. Bulletin of the Atomic Scientists, 48(7):35, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberg:1992:DFR

[GP92b] Stanley Goldberg and Thomas Powers. Declassified files reopen “Nazi bomb” debate. Bulletin of the Atomic Scientists, 48(7):32– 35, 38–40, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DR92].

Goldberg:1992:GFH

[GP92c] Stanley Goldberg and Thomas Powers. The “guests” at Farm Hall. Bulletin of the Atomic Scientists, 48(7):36–37, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goodby:2018:NBP

[GP18] James E. Goodby and Theodore A. Postol. A new boost-phase missile defense system — and its diplomatic uses in the North Korea dispute. Bulletin of the Atomic Scientists, 74(4):210–219, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1486578.

Gomer:1981:SRJ

[GPR81] Robert Gomer, John W. Powell, and Bert V. A. Röling. Special report: Japan’s biological weapons: 1930–1945. Bulletin of the Atomic Scientists, 37(8):43, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See update [Ano82d].

Green:1962:FGP

[GR62] Harold P. Green and Alan Rosenthal. Fusion of government power. Bulletin of the Atomic Scientists, 18(6):12–16, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Golovin:1967:FSB

[GR67] Igor Golovin and Eugene Rabinowitch. Father of the Soviet bomb. Bulletin of the Atomic Scientists, 23(10):11–18, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Garwin:1979:BMP

[GR79] Richard L. Garwin and Harry Reasoner. 60 Minutes on particle beam weapons. Bulletin of the Atomic Scientists, 35(2):50– 52, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter to CBS producer Don Hewitt with unedited script of film interview of Garwin by Reasoner.

Gray:1960:CW

[Gra60] J. Glenn Gray. The chances of war. Bulletin of the Atomic Scientists, 16(8):306–310, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Coh60].

Gray:1961:LEP

[Gra61] William Gray. Letter to the Editor: Problem of identity. Bulletin of the Atomic Scientists, 17(9):393, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab61b].

Graebner:1962:BBC

[Gra62] Norman A. Graebner. Books: The Cold War and Its Origins, by D. F. Fleming. Bulletin of the Atomic Scientists, 18(6):39–40, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1966:AP

[Gra66] Jack Gray. Agrarian policies. Bulletin of the Atomic Scientists, 22 (6):32–39, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1969:EM

[Gra69] Jack Gray. The economics of Maoism. Bulletin of the Atomic Scientists, 25(2):42–51, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Gray:1973:SSA

[Gra73] Colin S. Gray. Social science and the arms race. Bulletin of the Atomic Scientists, 29(6):23–26, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gomer:1976:LWT

[GRA^{+}76a] Robert Gomer, Stuart A. Rice, Aaron Adler, R. S. Berry, E. H. McLaren, D. H. Miller, Jr., Ezra Sensibar, W. Swartz, and Hans Zeisel. Letter: We take issue. Bulletin of the Atomic Scientists, 32 (9):5, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h]. The authors, who are members of the Bulletin’s Board of Directors, express disagreement with the editor.

Graham:1976:LWD

[Gra76b] J. Cozzy Graham. Letter: Wealth and debt. Bulletin of the Atomic Scientists, 32(10):5–6, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ben76a].

Gray:1981:CCS

[Gra81a] Colin S. Gray. Commentary: Chacun à son gout [each according to his own taste]. Bulletin of the Atomic Scientists, 37(6):65, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Buc81].

Gray:1981:CIN

[Gra81b] Colin S. Gray. Commentary: Issues and non-issues in the nuclear policy debate. Bulletin of the Atomic Scientists, 37(10):47–49, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DJE^{+}82].

Graham:1982:SUS

[Gra82a] Loren R. Graham. The Soviet Union: Science in the Brezhnev era. Bulletin of the Atomic Scientists, 38(2):23–28, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano82k].

Grant:1982:REN

[Gra82b] Robert Grant. Rabinowitch Essay: Necessary negotiations. Bulletin of the Atomic Scientists, 38(3):18–21, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Graham:1983:SES

[Gra83] Loren R. Graham. Scientific exchanges with the Soviet Union. Bulletin of the Atomic Scientists, 39(5):2–3, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1984:CDD

[Gra84a] Colin S. Gray. Commentary: In defense of disarmament. Bulletin of the Atomic Scientists, 40(6):46–47, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pan84, Smi84c].

Grayson:1984:BNB

[Gra84b] Lisa Grayson. Book note: Action for Arms Control: A Guide for Individuals and Organizations , by the Nuclear Weapons Discussion Project of Priorities. Bulletin of the Atomic Scientists, 40(6): 40, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grant:1985:ASE

[Gra85] James P. Grant. Africa: Silent emergency, loud aid. Bulletin of the Atomic Scientists, 41(8):46–49, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Graham:1986:PSH

[Gra86] Loren R. Graham. Perspective: Scientists, human rights, and the Soviet Union. Bulletin of the Atomic Scientists, 42(4):8–9, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [vL86, Hai86, Bat86].

Gray:1989:ACC

[Gra89] Colin S. Gray. Arms control criteria: People, not weapons, make war. Bulletin of the Atomic Scientists, 45(4):34, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1993:BRM

[Gra93] Peter Gray. Book review: The mythical Hanford: On the Home Front: The Cold War Legacy of the Hanford Nuclear Site, by Michele Stenehjem Gerber. Bulletin of the Atomic Scientists, 49 (3):45–46, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gray:1994:RSS

[Gra94] Peter Gray. Report: Stockpile stewardship: O’Leary v. Deutch. Bulletin of the Atomic Scientists, 50(6):10–12, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Graham:1998:SHT

[Gra98] Thomas Graham, Jr. Sanctions: Hang tough. Bulletin of the Atomic Scientists, 54(5):24–25, September/October 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Graves:2016:PPM

[Gra16] Alisha Graves. Population’s part in mitigating climate change: A US response. Bulletin of the Atomic Scientists, 72(3):186– 188, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1947:LED

[Gre47] John C. Green. Letter to the Editor: Distribution of declassified documents. Bulletin of the Atomic Scientists, 3(1):31, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1954:BBN

[Gre54] Howard Green. Books: The Next Million Years, by Charles G. Darwin. Bulletin of the Atomic Scientists, 10(2):55, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1955:BBA

[Gre55a] Harold Green. Books: The Atom Spy Hoax, by William Reuben. Bulletin of the Atomic Scientists, 11(8):309–310, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1955:USS

[Gre55b] Harold P. Green. The unsystematic security system. Bulletin of the Atomic Scientists, 11(4):118–122, 164, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Green:1958:BBS

[Gre58] Harold P. Green. Books: Loyalty and Security Employment Tests in the United States, by Ralph S. Brown, Jr. Bulletin of the Atomic Scientists, 14(7):276–277, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greene:1960:LI

[Gre60] Jerry Greene. Life itself? Bulletin of the Atomic Scientists, 16 (4):131, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Two-paragraph excerpt from Air Force Magazine, the January 1960, page 97.

Green:1962:NRA

[Gre62] Harold P. Green. News & reports: Atomic energy: Commission or administrator? economic impact of arms control. Bulletin of the Atomic Scientists, 18(10):33–38, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1963:AOD

[Gre63a] Philip Green. Alternatives to overkill: Dream and reality. Bulletin of the Atomic Scientists, 19(9):23–26, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greenwald:1963:LEL

[Gre63b] Eleanor Greenwald. Letter to the Editor: Law and mankind. Bulletin of the Atomic Scientists, 19(5):35, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1964:BBG

[Gre64a] Harold P. Green. Books: The Greatest Plot in History: How the Reds Stole the A-Bomb, by Ralph de Toledano. Bulletin of the Atomic Scientists, 20(2):29–30, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1964:QCD

[Gre64b] Harold P. Green. Q-clearance: The development of a personnel security program. Bulletin of the Atomic Scientists, 20(5):9–15, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1967:APT

[Gre67a] Harold P. Green. The AEC proposals — a threat to scientific freedom. Bulletin of the Atomic Scientists, 23(8):15–17, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1967:BBJ

[Gre67b] Harold P. Green. Books: The Journals of David E. Lilienthal. Volume I, The TVA Years, 1939–1945; Volume II, The Atomic Energy Years, 1945–1950; Volume III, Venturesome Years, 1950– 1955. Bulletin of the Atomic Scientists, 23(5):55–57, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1967:NTE

[Gre67c] Harold P. Green. The new technological era: A view from the law. Bulletin of the Atomic Scientists, 23(9):12–18, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greenberg:1967:CTS

[Gre67d] Daniel S. Greenberg. Comments: It’s time for science to act its political age. Bulletin of the Atomic Scientists, 23(8):36–37, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greenberg:1967:CSV

[Gre67e] David F. Greenberg. Comments: Students’ voices. Bulletin of the Atomic Scientists, 23(3):29–30, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1968:AIC

[Gre68a] Harold P. Green. AEC information control regulations. Bulletin of the Atomic Scientists, 24(5):41–43, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1968:BBW

[Gre68b] Murray Green. Books: The Wind and Beyond: Theodore von Karman, by Theodore von Karman with Lee Edson. Bulletin of the Atomic Scientists, 24(7):29, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1971:BBS

[Gre71] Harold P. Green. Books: Atomic Shield 1947/1952, by Richard G. Hewlett and Francis Duncan. Bulletin of the Atomic Scientists, 27(1):48–50, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greenbaum:1972:CNL

[Gre72] Leonard Greenbaum. Communications: The national laboratories. Bulletin of the Atomic Scientists, 28(9):2, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sac72b].

Greeley:1973:RPP

[Gre73a] Andrew M. Greeley. Reflections on politics and pluralism: A response to Jeremiads. Bulletin of the Atomic Scientists, 29(5): 31–34, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Obe74].

Greeley:1973:SDO

[Gre73b] Andrew M. Greeley. A scrapyard for the Daley organization? Bulletin of the Atomic Scientists, 29(2):9–14, February 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greeley:1973:WDA

[Gre73c] Andrew M. Greeley. A word of dissent: America first revisited. Bulletin of the Atomic Scientists, 29(9):50–51, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1977:OCS

[Gre77] Harold Green. The Oppenheimer case: a study in the abuse of law. Bulletin of the Atomic Scientists, 33(7):12–16, 56–61, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1978:RDC

[Gre78] Harold P. Green. The recombinant DNA controversy: a model of public influence. Bulletin of the Atomic Scientists, 34(9):12– 16, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bal78, DG79].

Green:1980:KC

[Gre80] Harold P. Green. On the Kemeny Commission. Bulletin of the Atomic Scientists, 36(3):46–48, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1981:BCA

[Gre81] Harold P. Green. ‘Born classified’ in the AEC: A legal perspective. Bulletin of the Atomic Scientists, 37(10):28–30, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1982:APP

[Gre82a] Harold P. Green. Atoms for peace: The peculiar politics of nuclear power. Bulletin of the Atomic Scientists, 38(10):59–65, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1982:BRB

[Gre82b] Harold P. Green. Book reviews: Born Secret, by A. DeVolpi, G. E. Marsh, T. A. Postol, and G. S. Stanford; The Secret That Exploded, by Howard Morland. Bulletin of the Atomic Scientists, 38 (5):51–53, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [WG82].

Green:1983:RDB

[Gre83a] Bill Green. Reagan’s defense budget: Mortgaging our future. Bulletin of the Atomic Scientists, 39(4):28–29, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gregg:1983:CAS

[Gre83b] Charles T. Gregg. Commentary: Anthrax in Sverdlovsk? Bulletin of the Atomic Scientists, 39(8):63, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zil83a] and comment [Zil83b].

Greb:1984:SWP

[Gre84a] G. Allen Greb. A system-wide program. Bulletin of the Atomic Scientists, 40(10):26s–27s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:1984:CC

[Gre84b] Leon Morton Green. Commentary: Congratulations. Bulletin of the Atomic Scientists, 40(7):56, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lig84, Bau84].

Greenhouse:1987:BRB

[Gre87] Carol J. Greenhouse. Book review: Blessed Assurance: At Home with the Bomb in Amarillo, Texas, by A. G. Mojtabai. Bulletin of the Atomic Scientists, 43(3):47–48, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Greenough:2005:ACS

[Gre05] William B. Greenough III. Article commentary: Security for whose sake? Bulletin of the Atomic Scientists, 61(5):22–23, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/5/22.

Greene:2007:SMP

[Gre07] Michael Greene. The science of Mideast peace. Bulletin of the Atomic Scientists, 63(4):80, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/4/80.

Greising:2008:CF

[Gre08] David Greising. The carbon frontier. Bulletin of the Atomic Scientists, 64(3):32–37, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/3/32.

Greenland:2012:UEW

[Gre12] Sander Greenland. Underestimating effects: Why causation probabilities need to be replaced in regulation, policy, and the law. Bulletin of the Atomic Scientists, 68(3):76–83, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/3/76.

Grego:2018:UGB

[Gre18] Laura Grego. US ground-based midcourse missile defense: Expensive and unreliable. Bulletin of the Atomic Scientists, 74 (4):220–226, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1486592. See corrigendum

[Ano18].

Green:2023:OCS

[Gre23] Harold P. Green. The Oppenheimer case: a study in the abuse of law. Bulletin of the Atomic Scientists, 79(4):255–264, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2223085.

Griswold:1965:WPC

[Gri65] S. Smith Griswold. What pollution costs. Bulletin of the Atomic Scientists, 21(6):12–16, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Griffin:1983:BNB

[Gri83a] Michael Griffin. Book note: London After the Bomb, by Owen Greene, Barry Rubin, Neil Turok, Philip Webber, and Graeme Wilkinson. Bulletin of the Atomic Scientists, 39(7):51, August/ September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grimmer:1983:CSM

[Gri83b] Derrick P. Grimmer. Commentary: A scientists’ moratorium. Bulletin of the Atomic Scientists, 39(2):54, February 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grinspoon:1984:CB

[Gri84] Lester Grinspoon. Crisis behavior. Bulletin of the Atomic Scientists, 40(4):25–28, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grimsson:1985:NAN

[Gri85] Olafur Ragnar Grimsson. Nuclear allergy: Nordic nuclear-free options. Bulletin of the Atomic Scientists, 41(6):25–28, June/ July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Griffiths:1987:NTK

[Gri87a] Franklyn Griffiths. “New thinking” in the Kremlin. Bulletin of the Atomic Scientists, 43(3):20–24, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Griswold:1987:LFN

[Gri87b] Arthur R. Griswold, Jr. Letter: Fix nuclear power. Bulletin of the Atomic Scientists, 43(9):54, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fla87] and comment [Har88b].

Grier:1991:PPP

[Gri91] Peter Grier. Poking and prying for peace. Bulletin of the Atomic Scientists, 47(10):25–29, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gromyko:1946:RPI

[Gro46] Andrej Gromyko. The Russian proposal for international control. Bulletin of the Atomic Scientists, 2(1–2):8–10, July 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gromyko:1947:GRC

[Gro47a] Andrej A. Gromyko. Gromyko replies to Cadogan questionnaire. Bulletin of the Atomic Scientists, 3(10):303–304, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gromyko:1947:RDR

[Gro47b] Andrej A. Gromyko. The Russian delegate’s reply to Osborn. Bulletin of the Atomic Scientists, 3(10):301–302, October 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gromyko:1947:SPA

[Gro47c] Andrej A. Gromyko. Soviet proposals for atomic energy control. Bulletin of the Atomic Scientists, 3(8):219–220, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gromyko:1948:DSC

[Gro48] Andrei A. Gromyko. A defense of the Soviet control plan. Bulletin of the Atomic Scientists, 4(6):191–192, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Groves:1949:AER

[Gro49] Leslie Groves. The atomic explosion in Russia: General Groves’ statement. Bulletin of the Atomic Scientists, 5(10):267, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1950:SOF

[Gro50] Clifford Grobstein. The scientists’ organization in 1949: The Federation of American Scientists. Bulletin of the Atomic Scientists, 6(2):58, 61, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grodzins:1951:BNL

[Gro51] Morton Grodzins. The basis of national loyalty. Bulletin of the Atomic Scientists, 7(12):356–362, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1952:NSF

[Gro52] Clifford Grobstein. National Science Foundation: Another round. Bulletin of the Atomic Scientists, 8(4):98–100, April 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1953:EFR

[Gro53a] Clifford Grobstein. Economy and Federal research. Bulletin of the Atomic Scientists, 9(6):221–222, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1953:FRD

[Gro53b] Clifford Grobstein. Federal research and development: Prospects, 1954. Bulletin of the Atomic Scientists, 9(8):299–304, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1953:NSF

[Gro53c] Clifford Grobstein. National Science Foundation: Report and appraisal. Bulletin of the Atomic Scientists, 9(2):39–42, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1953:WN

[Gro53d] Clifford Grobstein. Washington notes. Bulletin of the Atomic Scientists, 9(10):375, 389, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1954:WLPa

[Gro54a] Clifford Grobstein. Washington listening post. Bulletin of the Atomic Scientists, 10(1):26–27, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1954:WLPc

[Gro54b] Clifford Grobstein. Washington listening post. Bulletin of the Atomic Scientists, 10(3):103–104, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1954:WLPd

[Gro54c] Clifford Grobstein. Washington listening post. Bulletin of the Atomic Scientists, 10(4):136–137, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1954:WLPe

[Gro54d] Clifford Grobstein. Washington listening post. Bulletin of the Atomic Scientists, 10(5):169–170, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1956:PIS

[Gro56a] Clifford Grobstein. Part III: Science and social responsibility: The social conscience of U.S. science: Sketch of a decade. Bulletin of the Atomic Scientists, 12(7):241–246, September 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grodzins:1956:BBO

[Gro56b] Morton Grodzins. Books: The Oppenheimer Case: The Trial of a Security System, by Charles P. Curtis. Bulletin of the Atomic Scientists, 12(9):349, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SG69a, SG71].

Gromyko:1963:GU

[Gro63] Andrei Gromyko. Gromyko at the U.N. Bulletin of the Atomic Scientists, 19(9):43–45, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Groves:1970:SRJ

[Gro70] General Leslie Groves. Some recollections of July 16, 1945. Bulletin of the Atomic Scientists, 26(6):26–27, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1978:BRB

[Gro78a] Clifford Grobstein. Book reviews: A Sense of the Future, by J. Bronowski. Bulletin of the Atomic Scientists, 34(10):45–46, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grodzins:1978:LF

[Gro78b] Ruth Grodzins. Laura Fermi, 1907–1977. Bulletin of the Atomic Scientists, 34(5):2–3, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

GB:1978:DNP

[Gro78c] Groupe de Bellerive. Depolarizing nuclear power. Bulletin of the Atomic Scientists, 34(1):6–7, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobstein:1981:PCP

[Gro81] Clifford Grobstein. Perspective: From chance to purpose. Bulletin of the Atomic Scientists, 37(6):4–6, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Groer:1982:CHL

[Gro82] Peter G. Groer. Commentary: Hazards of low-level radiation. Bulletin of the Atomic Scientists, 38(2):46–47, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot81a].

Grossman:1983:NNF

[Gro83] Jerome Grossman. [Note on No first use: a view from Europe]. Bulletin of the Atomic Scientists, 39(3):22, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grossman:1984:CTD

[Gro84a] Jerome Grossman. Commentary: That’s not defense. Bulletin of the Atomic Scientists, 40(5):62–63, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa84a].

Grover:1984:BES

[Gro84b] Herbert D. Grover. Biological and ecological sciences. Bulletin of the Atomic Scientists, 40(10):9s–10s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grove:1986:ANF

[Gro86] Eric J. Grove. Allied nuclear forces complicate negotiations. Bulletin of the Atomic Scientists, 42(6):18–23, June/July 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gap86].

Gronlund:1998:AJK

[Gro98] Lisbeth Gronlund. ABM: Just kicking the can. Bulletin of the Atomic Scientists, 54(1):15–16, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gronlund:2001:RRR

[Gro01] Lisbeth Gronlund. Review, revamp, rethink NMD. Bulletin of the Atomic Scientists, 57(1):32–34, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/32.

Gronlund:2005:FAR

[Gro05] Lisbeth Gronlund. Fire, aim, ready. Bulletin of the Atomic Scientists, 61(5):67–68, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/5/67.

Grotto:2006:IHP

[Gro06a] Andrew J. Grotto. Iraq’s hot properties. Bulletin of the Atomic Scientists, 62(1):61–63, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/61.

Grotto:2006:TMP

[Gro06b] Andrew J. Grotto. A too-modest proposal. Bulletin of the Atomic Scientists, 62(3):54–56, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/3/54.

Grotto:2007:RCB

[Gro07] Andrew J. Grotto. Rotten carrots, broken sticks. Bulletin of the Atomic Scientists, 63(4):68–71, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/68.

Gronvall:2015:BOT

[Gro15] Gigi Kwik Gronvall. Biosecurity: The opportunities and threats of industrialization and personalization. Bulletin of the Atomic Scientists, 71(6):39–44, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grubel:1970:FSU

[Gru70a] Herbert G. Grubel. Foreign scientists in the United States. Bulletin of the Atomic Scientists, 26(4):9–12, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gruner:1970:BBC

[Gru70b] Wayne R. Gruner. Books: Scene of Change: A Lifetime in American Science, by Warren Weaver. Bulletin of the Atomic Scientists, 26(7):45, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gruener:1978:LSS

[Gru78] Garrett Gruener. Letter: Subsidizing solar. Bulletin of the Atomic Scientists, 34(1):2–3, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gruen:2004:TT

[Gru04] Madeleine Gruen. Talking to terrorists. Bulletin of the Atomic Scientists, 60(2):80–81, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/2/80.

Anonymous:1951:SOF

[GS51] Clifford Grobstein and A. H. Shapley. The scientists’ organizations in 1950: The Federation of American Scientists. Bulletin of the Atomic Scientists, 7(1):23–25, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grobman:1968:MSE

[GS68] Arnold B. Grobman and Laurence E. Strong. More on science education. Bulletin of the Atomic Scientists, 24(7):37–39, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bas68].

Glade:1981:CAF

[GS81] William Glade and James H. Street. Commentary: On academic freedom in Latin America. Bulletin of the Atomic Scientists, 37 (7):57–58, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Str81].

Gordon:1996:RBP

[GS96] Danielle Gordon and Lauren Spain. Reading between the party lines. Bulletin of the Atomic Scientists, 52(6):5–8, November/ December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gourley:2009:NED

[GS09] Bernard Gourley and Adam N. Stulberg. Nuclear energy development: Assessing aspirant countries. Bulletin of the Atomic Scientists, 65(6):20–29, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/65/6/20.

Garwin:1983:PBS

[GSB^{+}83] Richard Garwin, Carl Sagan, Hans A. Bethe, E. Margaret Burbidge, Clark R. Chapman, Thomas M. Donahue, Sidney D. Drell, Lee A. DuBridge, Herbert Freedman, Admiral Noel Gayler, Donald M. Hunten, Christopher C. Kraft, Vice-Admiral John Marshall Lee, Franklin A. Long, Carson Mark, James S. Martin, Jr., David Morrison, Philip Morrison, Bruce C. Murray, Gerry Neugebauer, Tobias Owen, Wolfgang K. H. Panofsky, Gordon H. Pettengill, William Pickering, Edward M. Purcell, I. I. Rabi, George W. Rathjens, Glenn T. Seaborg, Eugene M. Shoemaker, John A. Simpson, Edward C. Stone, James Van Allen, Victor F.

Weisskopf, Jerome B. Wiesner, Robert R. Wilson, Herbert F. York, et al. Perspective: Ban space weapons. Bulletin of the Atomic Scientists, 39(9):2–3, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ghamari-Tabrizi:2006:LF

[GT06] Sharon Ghamari-Tabrizi. Lethal fantasies. Bulletin of the Atomic Scientists, 62(1):20–22, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/20.

Guarnieri:2014:ETE

[GT14] Franck Guarnieri and Sébastien Travadel. Engineering thinking in emergency situations: a new nuclear safety concept. Bulletin of the Atomic Scientists, 70(6):79–86, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grin:1988:SD

[GU88] John Grin and Lutz Unterseher. The spiderweb defense. Bulletin of the Atomic Scientists, 44(7):28–30, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gubrud:2014:SKR

[Gub14] Mark Gubrud. Stopping killer robots. Bulletin of the Atomic Scientists, 70(1):32–42, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/70/1/32.full.pdf+html.

Gubrud:2015:GTF

[Gub15] Mark Gubrud. Going too fast: Time to ban hypersonic missile tests? : a US response. Bulletin of the Atomic Scientists, 71 (5):1–4, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/1.full.pdf+html.

Gueron:1966:RIS

[Gué66] Henri M. Guéron. Reflections on international security. Bulletin of the Atomic Scientists, 22(7):32–33, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gueron:1967:LLE

[Gué67] Jules Guéron. The lessons to be learned from Euratom. Bulletin of the Atomic Scientists, 23(3):38–41, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gueron:1969:LSP

[Gué69] Jules Guéron. The lack of scientific planning in Europe. Bulletin of the Atomic Scientists, 25(8):10–14, 25, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gueron:1970:AEC

[Gué70] Jules Guéron. Atomic energy in Continental Western Europe. Bulletin of the Atomic Scientists, 26(6):62–68, 116, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guertner:1979:CSM

[Gue79] Gary L. Guertner. Carter’s SALT: MAD or SAFE? Bulletin of the Atomic Scientists, 35(8):28–33, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guertner:1980:FRB

[Gue80] Gary L. Guertner. Film review: Protection in the Nuclear Age. Bulletin of the Atomic Scientists, 36(6):55–57, June 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guerra:2014:BBM

[Gue14] Héctor Guerra. Ban the bomb?: a Mexican response. Bulletin of the Atomic Scientists, 70(6):18–21, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guggenheim:1952:SBE

[Gug52] E. A. Guggenheim. Some British experiences. Bulletin of the Atomic Scientists, 8(7):231–232, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Guhl:1965:SM

[Guh65] A. M. Guhl. Sociobiology and man. Bulletin of the Atomic Scientists, 21(8):22–24, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sco66].

Guillain:1965:EOT

[Gui65] Robert Guillain. The explosion of October 16: Ten years of secrecy. Bulletin of the Atomic Scientists, 21(2):24–25, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guillemin:1977:LFD

[Gui77] Bernard Guillemin. Letter: Fyodorov’s dream. Bulletin of the Atomic Scientists, 33(1):3, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guillemin:2007:SS

[Gui07] Jeanne Guillemin. Seduced by the state. Bulletin of the Atomic Scientists, 63(5):14–16, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/5/14.

Gulbransen:1975:LSE

[Gul75] Earl A. Gulbransen. Letters: Not safe enough. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75, Pal76].

Gundersheim:1966:BRG

[Gun66] Art Gundersheim. Books: A reader’s guide to publications on China. Bulletin of the Atomic Scientists, 22(6):88–90, 92, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gunter:2004:NWS

[Gun04] Linda Clare Gunter. Nuclear waste: Showdown at Scanzano. Bulletin of the Atomic Scientists, 60(2):12–13, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/2/12.

Gunningham:2020:FLC

[Gun20] Neil Gunningham. Financing a low-carbon revolution. Bulletin of the Atomic Scientists, 76(5):228–232, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1806573.

Gurney:1975:AM

[Gur75] Ramsdell Gurney, Jr. Arms and the men. Bulletin of the Atomic Scientists, 31(10):23–33, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gurman:1988:CAW

[Gur88a] Lew Gurman. Choices and actions: Walking away from Star Wars. Bulletin of the Atomic Scientists, 44(3):6–8, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gurman:1988:DS

[Gur88b] Lew Gurman. Debunking SDI. Bulletin of the Atomic Scientists, 44(3):7, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See error report [Par88a].

Gurman:1988:DSW

[Gur88c] Lew Gurman. In defense of Star Wars. Bulletin of the Atomic Scientists, 44(3):8, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gushchenkov:1969:CPW

[Gus69] Ye. M. Gushchenkov. Can the Pacific warm the Arctic? Bulletin of the Atomic Scientists, 25(5):49, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gustafson:1970:RNP

[Gus70] Philip F. Gustafson. Reports: Nuclear power and thermal pollution: Zion, Illinois. Bulletin of the Atomic Scientists, 26(3): 17–23, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gustafson:1982:FC

[Gus82] Thane Gustafson. Further comment. Bulletin of the Atomic Scientists, 38(3):4–5, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pai82a].

Gustafson:1983:BRB

[Gus83] James M. Gustafson. Book reviews: The Nuclear Delusion: Soviet–American Relations in the Atomic Age, by George F. Kennan; Beyond the Cold War: A New Approach to the Arms Race and Nuclear Annihilation, by E. P. Thompson. Bulletin of the Atomic Scientists, 39(8):35–36, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gustafson:1985:BRB

[Gus85] James M. Gustafson. Book review: The Imperative of Responsibility: In Search of an Ethics for the Technological Age, by Hans Jonas. Bulletin of the Atomic Scientists, 41(4):58–59, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gusterson:1995:NTE

[Gus95] Hugh Gusterson. NIF-ty exercise machine. Bulletin of the Atomic Scientists, 51(5):22–26, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [CB95] and comment [Kat95].

Gusterson:1999:ASU

[Gus99] Hugh Gusterson. Los Alamos: Summer under siege. Bulletin of the Atomic Scientists, 55(6):36–41, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/6/36.

Gusterson:2001:TTD

[Gus01] Hugh Gusterson. Tall tales and deceptive discourses. Bulletin of the Atomic Scientists, 57(6):65–68, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/6/65.

Gusterson:2004:PBP

[Gus04] Hugh Gusterson. People of the bomb: portraits of America’s nuclear complex. University of Minnesota Press, Minneapolis, MN, USA, 2004. ISBN 0-8166-3859-4 (hardcover), 0-8166-3860-8 (paperback). xxvi + 312 pp. LCCN U264.3 .G87 2004. URL http:

//www.loc.gov/catdir/toc/ecip0415/2004003476.html.

Gusterson:2007:TRP

[Gus07] Hugh Gusterson. Taking RRW personally. Bulletin of the Atomic Scientists, 63(4):42–45, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/4/42.

Gusterson:2009:NA

[Gus09] Hugh Gusterson. Narrating abolition. Bulletin of the Atomic Scientists, 65(3):13–18, May 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/65/3/13.

Gusterson:2011:AAN

[Gus11] Hugh Gusterson. The assault on Los Alamos National Laboratory: a drama in three acts. Bulletin of the Atomic Scientists, 67(6):9–18, November/December 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/6/9.

Gutiontov:1993:ARV

[Gut93a] Pavel Gutiontov. Aleksandr Rutskoi, vice-king? Bulletin of the Atomic Scientists, 49(1):30, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gutiontov:1993:GBF

[Gut93b] Pavel Gutiontov. Gennadi Burbulis — the first to fall. Bulletin of the Atomic Scientists, 49(1):29, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gutiontov:1993:VP

[Gut93c] Pavel Gutiontov. In with the Volsky plan? Bulletin of the Atomic Scientists, 49(1):35, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gutiontov:1993:GP

[Gut93d] Pavel Gutiontov. Out with the Gaidar plan? Bulletin of the Atomic Scientists, 49(1):34, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gutiontov:1993:SRP

[Gut93e] Pavel Gutiontov. Suppressing a rising press. Bulletin of the Atomic Scientists, 49(1):39–41, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Guzman:1952:TGN

[Guz52] Leonardo Guzman. The treatment of good neighbors. Bulletin of the Atomic Scientists, 8(7):252, 258, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Garwin:2023:HAN

[GvH23] Richard L. Garwin and Frank N. von Hippel. How to avoid nuclear war with China. Bulletin of the Atomic Scientists, 79 (2):57–64, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2178161.

Greenwald:1965:LEW

[GW65] Eleanor Greenwald and Cecil R. Welte. Letter to the Editor: What role scientists? Bulletin of the Atomic Scientists, 21(9): 23, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab65g].

Gronlund:1992:BI

[GW92] Lisbeth Gronlund and David C. Wright. Building an ICBM. Bulletin of the Atomic Scientists, 48(2):36, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gronlund:1998:WTD

[GW98] Lisbeth Gronlund and David Wright. What they didn’t do. Bulletin of the Atomic Scientists, 54(6):46–51, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gretton-Watson:2002:EWW

[GW02] Paul Gretton-Watson. Energy: Wasted at the wellhead. Bulletin of the Atomic Scientists, 58(5):22–23, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/5/22.

Gyarmati:1990:HDK

[Gya90] Istvan Gyarmati. Hungarian diplomat: Keep Soviets on German soil. Bulletin of the Atomic Scientists, 46(5):10–11, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gyawali:2014:NRH

[Gya14] Dipak Gyawali. The Nepali response: How to energize women. Bulletin of the Atomic Scientists, 70(2):9–12, March 2014. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Gyorgy:1977:LRT

[Gyo77] Anna Gyorgy. Letter: The real terrorists. Bulletin of the Atomic Scientists, 33(4):2, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dow77].

H:1984:BNB

[H.84a] S. H. Book note: Uncertain Power: The Struggle for a National Energy Policy, by Dorothy S. Zinberg. Bulletin of the Atomic Scientists, 40(8):42, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

H:1984:BRB

[H.84b] S. H. Book review: 1984 National Directory of Audiovisual Resources on Nuclear War and the Arms Race, edited by Karen Sayer and John Dowling. Bulletin of the Atomic Scientists, 40(9): 53, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

H:1985:BNB

[H.85] S. H. Book note: The Nuclear Chain Reaction Forty Years Later, edited by Robert Sachs. Bulletin of the Atomic Scientists, 41(1): 53, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harry:1985:LSU

[HA85] R. J. S. Harry and David Albright. Letter: Superphénix and U.S.– Euratom agreements. Bulletin of the Atomic Scientists, 41(3): 62–64, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alb84] and comment [Low85].

Haaland:1973:CDO

[Haa73] Carsten M. Haaland. Communications: Difference of opinion. Bulletin of the Atomic Scientists, 29(1):3, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kra72].

Haaland:1977:LVT

[Haa77] Carsten M. Haaland. Letter: Victory through détente. Bulletin of the Atomic Scientists, 33(5):3, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Arb77].

Haaland:1978:LLL

[Haa78] Carsten M. Haaland. Letter: Last laugh. Bulletin of the Atomic Scientists, 34(8):64, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap78b].

Hackett:1981:FRB

[Hac81] Ellen Hackett. Film review: Mission Mind Control. Bulletin of the Atomic Scientists, 37(4):50, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hadamard:1952:SFE

[Had52] Jacques Hadamard. Some French experiences. Bulletin of the Atomic Scientists, 8(7):240, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Haddow:1956:PIS

[Had56] Alexander Haddow. Part III: Science and social responsibility: The scientist as citizen. Bulletin of the Atomic Scientists, 12(7): 247–252, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hadd:1977:LII

[Had77] John R. Hadd. Letter: Ideological impasse. Bulletin of the Atomic Scientists, 33(9):3–4, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafstad:1947:IPG

[Haf47] Lawrence R. Hafstad. An introduction to the problem of guided missiles. Bulletin of the Atomic Scientists, 3(11):341–342, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafstad:1949:APA

[Haf49] Lawrence R. Hafstad. Atomic power for aircraft. Bulletin of the Atomic Scientists, 5(11):309–312, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafstad:1950:BHR

[Haf50] Lawrence R. Hafstad. The breeder and the homogeneous reactor. Bulletin of the Atomic Scientists, 6(2):37, 49, February 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafstad:1951:RPA

[Haf51] Lawrence R. Hafstad. Reactor program of the Atomic Energy Commission. Bulletin of the Atomic Scientists, 7(4):109–114, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafstad:1954:SAI

[Haf54] Lawrence R. Hafstad. Speech at the Atomic Industrial Forum, September 28, 1954. Bulletin of the Atomic Scientists, 10(10): 392, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafner:1967:GS

[Haf67] E. M. Hafner. Galactic signals. Bulletin of the Atomic Scientists, 23(6):50–52, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafemeister:1979:BRB

[Haf79] David Hafemeister. Book review: Light Water, by Irvin C. Bupp and Jean-Claude Derian. Bulletin of the Atomic Scientists, 35(8): 38–39, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafemeister:1984:CRB

[Haf84] David Hafemeister. Commentary: A response to the build-down. Bulletin of the Atomic Scientists, 40(3):60–61, March 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafemeister:1985:AVT

[Haf85] David Hafemeister. Advances in verification technology. Bulletin of the Atomic Scientists, 41(1):35–40, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hafemeister:2005:SP

[Haf05] David Hafemeister. A secrecy primer. Bulletin of the Atomic Scientists, 61(3):23–25, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/23.

Haggerty:1977:BME

[Hag77] Robert J. Haggerty. Books: Is the medical establishment injurious to your health? Medical Nemesis: The Expropriation of Health, by Ivan Illich. Bulletin of the Atomic Scientists, 33(2):62–63, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hahn:1961:LER

[Hah61] Walter F. Hahn. Letter to the Editor: Review protested. Bulletin of the Atomic Scientists, 17(5–6):168, 255, May/June 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LM61b, LM61a].

Hahn:1967:OHAa

[Hah67a] Otto Hahn. Otto Hahn: Autobiographical notes [part 1 of 2]. Bulletin of the Atomic Scientists, 23(3):19–24, March 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from the 1966 book Otto Hahn: A Scientific Autobiography. See also part 2 [Hah67b].

Hahn:1967:OHAb

[Hah67b] Otto Hahn. Otto Hahn: Autobiographical notes [part 2 of 2]. Bulletin of the Atomic Scientists, 23(4):22–28, April 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from the 1966 book Otto Hahn: A Scientific Autobiography. See also part 1 [Hah67a].

Haissinsky:1954:LEM

[Hai54] M. Haissinsky. Letter to the Editor: Madame Joliot and the American Chemical Society. Bulletin of the Atomic Scientists, 10(2):56, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Eme54] and comment [Tod55].

Hailsham:1962:IIC

[Hai62] Lord Hailsham. The imperatives of international cooperation. Bulletin of the Atomic Scientists, 18(10):9–11, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haifley:1981:CDE

[Hai81] Dan Haifley. Commentary: Defining energy use. Bulletin of the Atomic Scientists, 37(2):64, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RW80].

Haines:1986:LSH

[Hai86] Andrew Haines. Letter: Soviet human rights. Bulletin of the Atomic Scientists, 42(7):68, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra86].

Holzman:1980:SRM

[HAIT80] Franklyn D. Holzman, Les Aspin, John David Isaacs, and Kosta Tsipis. Special report: The military sectors. Bulletin of the Atomic Scientists, 36(6):23, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hake:1962:IIF

[Hak62] David L. Hake. Improving the information flow. Bulletin of the Atomic Scientists, 18(9):21–22, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hall:1956:SP

[Hal56] Harry S. Hall. Scientists and politicians. Bulletin of the Atomic Scientists, 12(2):46–52, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halperin:1966:CNPa

[Hal66a] Morton H. Halperin. China and nuclear proliferation. I. Bulletin of the Atomic Scientists, 22(9):4–10, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halperin:1966:CNPb

[Hal66b] Morton H. Halperin. China and nuclear proliferation. II. Bulletin of the Atomic Scientists, 22(10):18–24, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halsted:1971:LAA

[Hal71] Thomas A. Halsted. Lobbying against the ABM, 1967–1970. Bulletin of the Atomic Scientists, 27(4):23–28, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hall:1975:LWH

[Hal75a] Charles A. S. Hall. Look what’s happening to our Earth: The biosphere, the industriosphere and their interactions. Bulletin of the Atomic Scientists, 31(3):11–21, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halsted:1975:SNW

[Hal75b] Thomas A. Halsted. The spread of nuclear weapons — is the dam about to burst? Non-Proliferation Treaty in deep trouble as world hovers on brink of a new arms race. Bulletin of the Atomic Scientists, 31(5):8–11, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hall:1976:LUS

[Hal76a] Syd M. Hall. Letter: Unheeded supplications. Bulletin of the Atomic Scientists, 32(2):3, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75g].

Hall:1976:LCD

[Hal76b] William Cornelius Hall. Letter: For civil defense. Bulletin of the Atomic Scientists, 32(1):3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pri75b].

Halperin:1983:EFA

[Hal83] Morton H. Halperin. Editorial: The freeze in arms control. Bulletin of the Atomic Scientists, 39(3):2–3, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halpern:1984:CWT

[Hal84] Jennifer Halpern. Commentary: Working with technology for peace. Bulletin of the Atomic Scientists, 40(1):61–62, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halperin:1985:SNS

[Hal85] Morton H. Halperin. Secrecy and national security. Bulletin of the Atomic Scientists, 41(7):114–117, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [KH85].

Halperin:1989:ACC

[Hal89] Morton H. Halperin. Arms control criteria: From primer to policy. Bulletin of the Atomic Scientists, 45(4):31–32, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hallin:1991:TCL

[Hal91] Daniel Hallin. TV’s clean little war. Bulletin of the Atomic Scientists, 47(4):16–19, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hallin:1993:BRV

[Hal93a] Daniel C. Hallin. Book reviews: Video games: Hotel Warriors: Covering the Gulf War, by John J. Fialka; The Persian Gulf TV War, by Douglas Kellner; War and the Media: Propaganda and Persuasion in the Gulf War, by Philip M. Taylor; War Game: L’information et la guerre, by Dominique Wolton. Bulletin of the Atomic Scientists, 49(7):52–54, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Halverson:1993:TTB

[Hal93b] Thomas Halverson. Ticking time bombs: East Bloc reactors. Bulletin of the Atomic Scientists, 49(6):43–48, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hales:2007:AP

[Hal07] Peter Bacon Hales. The atomic Psyche. Bulletin of the Atomic Scientists, 63(4):66–68, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/4/66.

Hammarskjold:1958:DNT

[Ham58] Dag Hammarskjold. [Disarmament and nuclear test detection]. Bulletin of the Atomic Scientists, 14(8):287, October 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from Introduction to Annual Report of Work of United Nations.

Hamermesh:1961:BBC

[Ham61a] Morton Hamermesh. Books: Soviet Cultural Diplomacy, by Frederick C. Barghoorn. Bulletin of the Atomic Scientists, 17(2):76– 77, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamermesh:1961:BBP

[Ham61b] Morton Hamermesh. Books: Soviet Policy Towards International Control of Atomic Energy, by Joseph L. Nogee. Bulletin of the Atomic Scientists, 17(10):437, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:1961:MRU

[Ham61c] Walter A. Hamilton. McKinney review of U.S. atomic policies. Bulletin of the Atomic Scientists, 17(1):25–29, 40, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kaz61].

Hamilton:1963:NRF

[Ham63] L. D. Hamilton. News & reports: Fallout and countermeasures. Bulletin of the Atomic Scientists, 19(7):36–40, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:1972:RS

[Ham72] L. D. Hamilton. On radiation standards. Bulletin of the Atomic Scientists, 28(3):30–33, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hammond:1976:VSS

[Ham76] George S. Hammond. The value system in the scientific subculture. Bulletin of the Atomic Scientists, 32(10):36–40, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dud77a].

Hammond:1983:BRB

[Ham83] Alexander Hammond. Book review: No Place to Hide, 1946/ 1984, by David Bradley. Bulletin of the Atomic Scientists, 39 (9):37–39, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:1990:PFS

[Ham90] Lee H. Hamilton. Perspective: Finish START now. Bulletin of the Atomic Scientists, 46(8):10, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:1991:DD

[Ham91a] John Maxwell Hamilton. From dissident to deputy. Bulletin of the Atomic Scientists, 47(5):16, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:1991:WPK

[Ham91b] John Maxwell Hamilton. Will pollution kill the revolution? Bulletin of the Atomic Scientists, 47(5):12–18, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [AG91].

Hamilton:1991:BRB

[Ham91c] Tom Hamilton. Book review: Beyond Bravo: Day of Two Suns: U.S. Nuclear Testing and the Pacific Islanders, by Jan Dibblin. Bulletin of the Atomic Scientists, 47(2):45, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:1993:BRN

[Ham93] John Maxwell Hamilton. Book review: Nuclear creep: Nonproliferation and the National Interest: America’s Response to the Spread of Nuclear Weapons, by Peter A. Clausen. Bulletin of the Atomic Scientists, 49(9):50–51, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamza:1998:ISS

[Ham98] Khidhir Hamza. Inside Saddam’s secret nuclear program. Bulletin of the Atomic Scientists, 54(5):26–33, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hamilton:2014:GPS

[Ham14] Clive Hamilton. Geoengineering and the politics of science. Bulletin of the Atomic Scientists, 70(3):17–26, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hancock:1946:UPC

[Han46] John Hancock. The US plan for control of atomic energy. Bulletin of the Atomic Scientists, 2(5–6):14–15, September 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Han:1966:RSC

[Han66] Suyin Han. Reflections on social change. Bulletin of the Atomic Scientists, 22(6):80–83, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Handler:1969:ORS

[Han69] Philip Handler. Other reactions in the science community: National Science Board. Bulletin of the Atomic Scientists, 25(6):47, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Handler:1975:BBB

[Han75] Paul Handler. Books: By Bread Alone, by Lester R. Brown with Erik P. Eckholm; Mankind at the Turning Point: The Second Report to the Club of Rome, by Mihajlo Mesarovic and Eduard Pestel. Bulletin of the Atomic Scientists, 31(6):47–49, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Handler:1976:DWB

[Han76] Philip Handler. Detlev Wulf Bronk 1897–1975. Bulletin of the Atomic Scientists, 32(2):6, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1979:CBC

[Han79] Chuck Hansen. Commentary: That Progressive case again. Bulletin of the Atomic Scientists, 35(8):59–60, October 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79i].

Handler:1987:ARS

[Han87a] Joshua Handler. Arms race at sea: Waging submarine warfare. Bulletin of the Atomic Scientists, 43(7):40–43, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hanson:1987:SIE

[Han87b] Philip Hanson. Soviet industrial espionage. Bulletin of the Atomic Scientists, 43(3):25–29, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1989:LMD

[Han89] Chuck Hansen. Letter: More dirty little secrets. Bulletin of the Atomic Scientists, 45(6):49–50, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gea89].

Handler:1990:BSC

[Han90a] Joshua Handler. Bulletin: Soviets confirm Baltic denuclearized. Bulletin of the Atomic Scientists, 46(10):5–6, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1990:MAD

[Han90b] Chuck Hansen. 1,000 more accidents declassified. Bulletin of the Atomic Scientists, 46(5):9, 41, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1990:BRO

[Han90c] Chuck Hansen. Book review: Oops! The Hidden Cost of Deterrence, by Chuck Hansen. Bulletin of the Atomic Scientists, 46(8): 42–43, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1990:LNT

[Han90d] Chuck Hansen. Letter: Neptune test collapse a 115-ton accident. Bulletin of the Atomic Scientists, 46(2):47, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [O’N89a].

Handler:1991:RSU

[Han91] Joshua Handler. Report: Soviet Union: The greening of Petropavlovsk. Bulletin of the Atomic Scientists, 47(8):11, 46, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Handler:1992:BSH

[Han92a] Joshua Handler. Bulletin: Send help, not charity. Bulletin of the Atomic Scientists, 48(4):6–7, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1992:BRD

[Han92b] Chuck Hansen. Book review: The dark side: Teller’s War: The Top-Secret Story Behind the Star Wars Deception, by William J. Broad. Bulletin of the Atomic Scientists, 48(4):45–46, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1992:LUS

[Han92c] Chuck Hansen. Letter: Ulam’s shock waves. Bulletin of the Atomic Scientists, 48(1):45, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BHM91].

Handler:1993:RNN

[Han93] Joshua Handler. Report: Nuclear navy: No sleep in the deep for Russian subs. Bulletin of the Atomic Scientists, 49(3):7–9, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1994:BRB

[Han94] Chuck Hansen. Book review: Nuclear Weapons Databook: British, French, and Chinese nuclear weapons, by Robert S. Norris, Andrew S. Burrows, and Richard W. Fieldhouse. Bulletin of the Atomic Scientists, 50(4):54–55, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1995:BRB

[Han95a] Chuck Hansen. Book review: The bomb, part 2: Dark Sun: The Making of the Hydrogen Bomb, by Richard Rhodes. Bulletin of the Atomic Scientists, 51(5):52–53, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://www.highbeam.com/doc/1G1-17290233.html.

Hansen:1995:LQ

[Han95b] Chuck Hansen. Letter: Without a “Q”. Bulletin of the Atomic Scientists, 51(6):73, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Han95c].

Hansen:1995:OSC

[Han95c] Chuck Hansen. Open secrets, closed minds. Bulletin of the Atomic Scientists, 51(4):16–17, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Han96b, Han95b].

Handler:1996:RAC

[Han96a] Joshua Handler. Report: Arms control: Russia ready for START III. Bulletin of the Atomic Scientists, 52(1):11–12, January/ February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1996:LUR

[Han96b] Chuck Hansen. Letter: Useful resource. Bulletin of the Atomic Scientists, 52(4):3, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Han95c].

Hansen:1997:LMF

[Han97] Chuck Hansen. Letter: More fallout on fallout. Bulletin of the Atomic Scientists, 53(5):3, 58, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM97].

Hansen:1998:BRA

[Han98a] Chuck Hansen. Book review: Not available in stores: U.S. Nuclear History: Nuclear Arms and Policy in the Missile Age, 1955– 1968, William Burr, project director, and Thomas S. Blanton and Malcolm Byrne, series editors. Bulletin of the Atomic Scientists, 54(6):60–61, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1998:LRP

[Han98b] Tom Hansen. Letter: Resume peace talks. Bulletin of the Atomic Scientists, 54(3):3, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nad98c].

Handler:2000:NAA

[Han00a] Joshua Handler. The Nikitin Affair: an acquittal at last? Bulletin of the Atomic Scientists, 56(2):17–19, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/2/17.

Hansen:2000:OL

[Han00b] Chuck Hansen. The oops list. Bulletin of the Atomic Scientists, 56(6):64–67, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/6/64.

Hansen:2001:BOS

[Han01] Chuck Hansen. Beware the old story. Bulletin of the Atomic Scientists, 57(2):52–55, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/52.

Handal:2002:AUF

[Han02] Nabil Handal. Aid under fire. Bulletin of the Atomic Scientists, 58(6):82–87, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/82.

Hansen:2005:CFF

[Han05] Keith Hansen. CTBT: Forecasting the future. Bulletin of the Atomic Scientists, 61(2):50–57, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/2/50.

Han:2017:CPR

[Han17] Hua Han. China’s proper role in the global nuclear order. Bulletin of the Atomic Scientists, 73(2):128–130, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1288451.

Haqqani:2005:PGD

[Haq05] Husain Haqqani. Pakistan’s great divide. Bulletin of the Atomic Scientists, 61(5):69–70, September/October 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/5/69.

Harris:1947:RRE

[Har47] Seymour E. Harris. Replies to Ridenour: An economist views the problem. Bulletin of the Atomic Scientists, 3(8):226–227, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartley:1950:FP

[Har50a] Livingston Hartley. False parallel. Bulletin of the Atomic Scientists, 6(6):172–173, 192, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartley:1950:MUE

[Har50b] Sir Harold Hartley. Man’s use of energy. Bulletin of the Atomic Scientists, 6(11):322–324, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hart:1954:RVM

[Har54] Hornell Hart. The remedies versus the menace. Bulletin of the Atomic Scientists, 10(6):197–205, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Spe54].

Harlow:1956:BBG

[Har56] James G. Harlow. Books: The Gifted Student as Future Scientist, by Paul E. Brandwein. Bulletin of the Atomic Scientists, 12(1): 24–25, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harrow:1959:LED

[Har59] Robert Harrow. Letter to the Editor: Disappointment in Bulletin. Bulletin of the Atomic Scientists, 15(10):433, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harding:1960:BBU

[Har60a] H. F. Harding. Books: The Uncertain Trumpet, by General Maxwell Taylor. Bulletin of the Atomic Scientists, 16(6):221, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartshorne:1960:LEM

[Har60b] Charles Hartshorne. Letter to the Editor: Meaning of freedom. Bulletin of the Atomic Scientists, 16(10):374, 423, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul60].

Harcourt:1961:LEP

[Har61] John Terry Harcourt. Letter to the Editor: Plan for Berlin. Bulletin of the Atomic Scientists, 17(3):121, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harvey:1962:LEC

[Har62] Cyrus I. Harvey, Jr. Letter to the Editor: Crisis decisions. Bulletin of the Atomic Scientists, 18(8):33, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harris:1964:BEP

[Har64a] Seymour E. Harris. Books: Economics and politics. Bulletin of the Atomic Scientists, 20(1):32–34, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartshorne:1964:LED

[Har64b] Charles Hartshorne. Letter to the Editor: Dobzhansky and psychicalistic views. Bulletin of the Atomic Scientists, 20(8):38, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dob64].

Hartshorne:1965:LES

[Har65] Charles Hartshorne. Letter to the Editor: Szilard and the realities of life. Bulletin of the Atomic Scientists, 21(5):35, May 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harris:1969:OMD

[Har69a] Richard Harris. One-man diplomacy? Bulletin of the Atomic Scientists, 25(2):60–65, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Harvey:1969:LLU

[Har69b] Mose L. Harvey. The lunar landing and the U.S.–Soviet equation. Bulletin of the Atomic Scientists, 25(7):28–32, 34–35, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1970:CTS

[Har70] Garrett Hardin. Comments: To trouble a star: The cost of intervention in nature. Bulletin of the Atomic Scientists, 26(1):

17–20, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CH71].

Harris:1971:CCE

[Har71] Robert G. Harris. Communications: On chemical escalation. Bulletin of the Atomic Scientists, 27(5):48, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1972:CLG

[Har72a] Garrett Hardin. Comment: Limits to growth — two views: We live on a spaceship. Bulletin of the Atomic Scientists, 28(9):22– 25, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wat73b].

Hardin:1972:PSE

[Har72b] Garrett Hardin. Population skeletons in the environmental closet. Bulletin of the Atomic Scientists, 28(6):37–41, June 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1976:LFB

[Har76a] Garrett Hardin. Living with the Faustian bargain. Bulletin of the Atomic Scientists, 32(9):25–29, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1976:TSN

[Har76b] Garrett Hardin. The trans-science of ‘necessity’. Bulletin of the Atomic Scientists, 32(9):24, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartman:1977:LRF

[Har77a] Stan Hartman. Letter: The reality of faith. Bulletin of the Atomic Scientists, 33(6):5, 8, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hoa77].

Harwood:1977:LPN

[Har77b] Paul D. Harwood. Letter: Proving a negative. Bulletin of the Atomic Scientists, 33(3):6, 8, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei76b].

Harnik:1979:CEE

[Har79a] Peter Harnik. Commentary: The ethics of energy production and use: debate within the National Council of Churches. Bulletin of the Atomic Scientists, 35(2):5–9, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [AH79].

Harris:1979:BRB

[Har79b] Neil Harris. Book review: Small Comforts for Hard Times. Humanists on Public Policy, edited by Michael Mooney and Florian Stuber. Bulletin of the Atomic Scientists, 35(5):36, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartman:1979:CRE

[Har79c] Hyman Hartman. Commentary: Reflections on Einstein’s 100th birthday. Bulletin of the Atomic Scientists, 35(10):9, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harte:1981:BRB

[Har81] John Harte. Book review: Extinction — The Causes and Consequences of the Disappearance of Species, by Paul and Anne Ehrlich. Bulletin of the Atomic Scientists, 37(8):37, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harvey:1983:CUR

[Har83] Clifford Harvey. Commentary: On ‘The Ultimate Resource’. Bulletin of the Atomic Scientists, 39(8):64, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SD83].

Hardin:1984:CPA

[Har84a] Russell Hardin. Contracts, promises and arms control. Bulletin of the Atomic Scientists, 40(8):14–17, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harriman:1984:PIR

[Har84b] W. Averell Harriman. Perspective: If the Reagan pattern continues …. Bulletin of the Atomic Scientists, 40(3):2, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harris:1984:CUI

[Har84c] J. Merle Harris. Commentary: Unilateral initiatives. Bulletin of the Atomic Scientists, 40(7):56, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardenbergh:1985:ON

[Har85a] Chalmers Hardenbergh. The other negotiations. Bulletin of the Atomic Scientists, 41(6):42–44, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1985:BRB

[Har85b] Russell Hardin. Book review: The Illogic of American Nuclear Strategy, by Robert Jervis. Bulletin of the Atomic Scientists, 41 (2):47–48, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1985:R

[Har85c] Russell Hardin. A rejoinder. Bulletin of the Atomic Scientists, 41 (4):53–54, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bil85].

Harbutt:1986:ICC

[Har86a] Fraser J. Harbutt. The iron curtain: Churchill, America, and the origins of the Cold War. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-503817-7, 0-19-505422-9 (paperback). xiv + 370 pp. LCCN DA588 .H27

1986. URL http://site.ebrary.com/lib/stanford/Doc?id=
10086780;   http://www.myilibrary.com?id=44044.

Hardenbergh:1986:ONa

[Har86b] Chalmers Hardenbergh. The other negotiations. Bulletin of the Atomic Scientists, 42(1):45–47, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardenbergh:1986:ONb

[Har86c] Chalmers Hardenbergh. The other negotiations. Bulletin of the Atomic Scientists, 42(6):42–44, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1986:SWP

[Har86d] William Hartung. Star Wars pork barrel. Bulletin of the Atomic Scientists, 42(1):20–24, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1986:BRB

[Har86e] William D. Hartung. Book review: The American Arms Supermarket, by Michael T. Klare. Bulletin of the Atomic Scientists, 42(6):56–57, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardenbergh:1987:ONa

[Har87a] Chalmers Hardenbergh. The other negotiations. Bulletin of the Atomic Scientists, 43(2):48–49, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardenbergh:1987:ONb

[Har87b] Chalmers Hardenbergh. The other negotiations. Bulletin of the Atomic Scientists, 43(7):52–53, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1987:MTW

[Har87c] William Hartung. Major Third World arms importers and their suppliers, 1981–1985. Bulletin of the Atomic Scientists, 43(10): 30, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1987:NVA

[Har87d] William Hartung. Nations vie for arms markets. Bulletin of the Atomic Scientists, 43(10):27–35, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1987:RRA

[Har87e] William Hartung. The Reagan revival of arms deals. Bulletin of the Atomic Scientists, 43(6):20–25, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardenbergh:1988:ON

[Har88a] Chalmers Hardenbergh. The other negotiations. Bulletin of the Atomic Scientists, 44(5):43–44, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hardin:1988:LLD

[Har88b] Garrett Hardin. Letter: “Longage” of demand? Bulletin of the Atomic Scientists, 44(1):54, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gri87b].

Hartung:1988:BRB

[Har88c] William Hartung. Book reviews: Veil: The Secret Wars of the CIA 1981–1987, by Bob Woodward; The Crimes of Patriots: A True Tale of Dope, Dirty Money, and the CIA, by Jonathan Kwitney. Bulletin of the Atomic Scientists, 44(6):48–49, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harvey:1988:DA

[Har88d] Hal Harvey. Defense without aggression. Bulletin of the Atomic Scientists, 44(7):12–15, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hart:1989:CFR

[Har89a] Kathleen Hart. Corporate-funded research may be hazardous to your health. Bulletin of the Atomic Scientists, 45(3):32, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hart:1989:AFB

[Har89b] Kathleen Hart. Is academic freedom bad for business? Bulletin of the Atomic Scientists, 45(3):28–34, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SH89a].

Hardt:1990:CC

[Har90a] John P. Hardt. Conversion or chaos? Bulletin of the Atomic Scientists, 46(1):20, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1990:KM

[Har90b] William D. Hartung. Korean missiles. Bulletin of the Atomic Scientists, 46(9):24, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1990:MA

[Har90c] William D. Hartung. The M-16 affair. Bulletin of the Atomic Scientists, 46(9):23, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1990:UKJ

[Har90d] William D. Hartung. U.S.–Korea jet deal boosts arms trade. Bulletin of the Atomic Scientists, 46(9):18–24, November 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LH91].

Hartung:1990:WB

[Har90e] William D. Hartung. Wooing the buyer. Bulletin of the Atomic Scientists, 46(9):21, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harris:1991:LER

[Har91a] Edward F. Harris. Letter: Electromagnetic radiation and VDTs. Bulletin of the Atomic Scientists, 47(1):46, January 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jau90a].

Hartung:1991:BAB

[Har91b] William Hartung. The boom at the arms bazaar. Bulletin of the Atomic Scientists, 47(8):14–20, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1992:BRGb

[Har92a] William D. Hartung. Book review: Gun man meets gunman: Bull’s Eye: The Assassination and Life of Supergun Inventor Gerald Bull, by James Adams. Bulletin of the Atomic Scientists, 48(8):42–43, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1992:BRGa

[Har92b] William D. Hartung. Book review: Guns as bread-and-butter: The Death Lobby: How the West Armed Iraq, by Kenneth R. Timmerman. Bulletin of the Atomic Scientists, 48(7):41–42, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hartung:1993:WUA

[Har93] William D. Hartung. Welcome to the U.S. arms superstore. Bulletin of the Atomic Scientists, 49(7):20–26, September 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harwit:1996:EDL

[Har96] Martin Harwit. An exhibit denied: lobbying the history of Enola Gay. Copernicus, New York, NY, USA, 1996. ISBN 0-387-94797-3 (hardcover). xxv + 477 pp. LCCN D767.25.H6 H348 1996.

Hartung:1999:BBP

[Har99] William D. Hartung. Billions for bombs; pennies for peacekeeping. Bulletin of the Atomic Scientists, 55(5):6, September/ October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
5/6.

Hartung:2001:SAW

[Har01] William D. Hartung. Stop arming the world. Bulletin of the Atomic Scientists, 57(1):34–36, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/34.

Hartung:2004:PW

[Har04] William D. Hartung. Privatizing war. Bulletin of the Atomic Scientists, 60(2):78–79, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/2/78.

Harris:2005:TKA

[Har05] Shane Harris. Terrorism: Knockoff artists. Bulletin of the Atomic Scientists, 61(5):16–18, September/October 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/5/16.

Harris:2006:BB

[Har06] Shane Harris. The bug bloggers. Bulletin of the Atomic Scientists, 62(1):38–43, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/1/38.

Harvey:2007:NNS

[Har07] John R. Harvey. Nonproliferation’s new soldier. Bulletin of the Atomic Scientists, 63(4):32–33, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/32.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hart:2014:URH

[Har14a] Corinne Hart. The US response: How to energize women. Bulletin of the Atomic Scientists, 70(2):17–20, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hart:2014:ZCW

[Har14b] John Hart. In the zone? Chemical weapons and the Middle East: The US response. Bulletin of the Atomic Scientists, 70(3):7–9, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harvey:2014:RUN

[Har14c] John R. Harvey. Reforming the US nuclear weapons enterprise. Bulletin of the Atomic Scientists, 70(6):29–38, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harrington:2022:SRO

[Har22] Brooke Harrington. Sanctioning Russia’s oligarchs — with shame. Bulletin of the Atomic Scientists, 78(6):329–333, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2132740.

Hasterlick:1957:BBE

[Has57] Robert J. Hasterlick. Books: Medical Effects of the Atomic Bomb in Japan, Ashley W. Oughterson and Shields Warren. Bulletin of the Atomic Scientists, 13(1):35, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haskins:1960:SSR

[Has60] Caryl P. Haskins. Society and scientific research. Bulletin of the Atomic Scientists, 16(5):146–150, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haselkorn:1967:SSP

[Has67] Robert Haselkorn. Science and space policy: Editorial introduction. Bulletin of the Atomic Scientists, 23(3):2–3, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hass:1968:COS

[Has68] Ernst Hass. Common opponent sought … and found? Bulletin of the Atomic Scientists, 24(9):8–11, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haselkorn:1971:BBM

[Has71] Robert Haselkorn. Books: Life on Man, by Theodor Rosebury. Bulletin of the Atomic Scientists, 27(1):51, January 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hastings:1985:LDC

[Has85] Tom Hastings. Letter: Don’t count on Reagan’s negotiators. Bulletin of the Atomic Scientists, 41(6):57, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haslam:1988:STF

[Has88] Jonathan Haslam. Soviets take fresh look at Western Europe. Bulletin of the Atomic Scientists, 44(4):38–42, May 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hatch:1985:LNN

[Hat85] Senator Orrin G. Hatch. Letter: NATO nuclear strategy. Bulletin of the Atomic Scientists, 41(6):53, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha85c].

Hauser:1964:MMM

[Hau64] Philip M. Hauser. Man and more men: The population prospects. Bulletin of the Atomic Scientists, 20(6):4–8, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hauser:1968:CCM

[Hau68] Philip Hauser. Current comments: Mounting chaos at home. Bulletin of the Atomic Scientists, 24(1):56–58, January 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hauser:1989:LAA

[Hau89] Jack C. Hauser. Letter: ACM the acme. Bulletin of the Atomic Scientists, 45(8):42–43, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark89d].

Hauer:2017:UCM

[Hau17] Jerome M. Hauer. US cities are not medically prepared for a nuclear detonation. Bulletin of the Atomic Scientists, 73(4): 215–219, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1338003.

Haviland:1969:RSB

[Hav69] Robert P. Haviland. Reports: Space broadcasting — how, when, and why. Bulletin of the Atomic Scientists, 25(3):39–42, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Havens:2003:TRB

[Hav03] Jerry Havens. Terrorism: Ready to blow? Bulletin of the Atomic Scientists, 59(4):16–18, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/4/16.

Havens:2004:LSS

[Hav04] Jerry Havens. LNG: Safety in science. Bulletin of the Atomic Scientists, 60(1):30–31, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/1/30.

Hawkins:1947:BBU

[Haw47] David Hawkins. Books: On Understanding Science: An Historical Approach, by James B. Conant. Bulletin of the Atomic Scientists, 3(10):311–312, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawkins:1950:BRB

[Haw50] David Hawkins. Book review: The History of Nature, by C. F. von Weizsäcker. Bulletin of the Atomic Scientists, 6(11):348, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawley:1951:UDD

[Haw51] Amos H. Hawley. Urban dispersal and defense. Bulletin of the Atomic Scientists, 7(10):307–312, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawkins:1956:BBH

[Haw56] David Hawkins. Books: Science in History, by J. D. Bernal. Bulletin of the Atomic Scientists, 12(2):60, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawkins:1958:BBH

[Haw58a] David Hawkins. Books: Science and Human Values, by J. Bronowski. Bulletin of the Atomic Scientists, 14(7):277, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawkins:1958:BBC

[Haw58b] David Hawkins. Books: The Common Sense of the Exact Sciences, by William Kingdon Clifford. Bulletin of the Atomic Scientists, 14(5):196, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawkins:1973:PAA

[Haw73] Gerald S. Hawkins. Prehistoric astronomy: Astro-archaeology — the unwritten evidence. Bulletin of the Atomic Scientists, 29(8): 58–64, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hawkings:1994:LOB

[Haw94] David Hawkings. Letter: Oppenheimer and Bohr. Bulletin of the Atomic Scientists, 50(5):60, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hayasaka:1955:OLA

[Hay55] Ichiro Hayasaka. Open letter: Appeal to Western evolutionists against the hydrogen bomb. Bulletin of the Atomic Scientists, 11 (1):34, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haybittle:1964:ES

[Hay64] John Haybittle. Ethics for the scientist. Bulletin of the Atomic Scientists, 20(5):23–24, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Daf64].

Hayes:1976:LKD

[Hay76] Peter Hayes. Letter: The killing dynamic. Bulletin of the Atomic Scientists, 32(8):2, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gal76, Dum76].

Hayes:1978:PTW

[Hay78] Denis Hayes. Priorities for the Third World. Bulletin of the Atomic Scientists, 34(6):9–10, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hayward:1979:HNW

[Hay79] J. T. Hayward. How a nuclear war would be fought … in the Middle East. Bulletin of the Atomic Scientists, 35(5):26–28, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hay:1981:CWN

[Hay81] Edward Hay. Commentary from Washington: NATO outspends Warsaw. Bulletin of the Atomic Scientists, 37(6):9, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hayes:1993:PWN

[Hay93] Peter Hayes. Perspective: What North Korea wants. Bulletin of the Atomic Scientists, 49(10):8–10, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hazzard:1982:CWR

[Haz82] Shirley Hazzard. Commentary: A write reflects. Bulletin of the Atomic Scientists, 38(3):43–44, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hagan:1963:MVM

[HB63] Roger Hagan and Bart Bernstein. Military value of missiles in Cuba. Bulletin of the Atomic Scientists, 19(2):8–13, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harris:1969:RGP

[HB69] Richard Harris and William Brugger. A reader’s guide to publications from and on China. Bulletin of the Atomic Scientists, 25(2):84–88, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Heyes:2010:SPG

[HB10] Alan J. Heyes and Wyn Q. Bowen. Silent partnership: The G-8’s nonproliferation program. Bulletin of the Atomic Scientists, 66(2):17–26, March 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/2/17.

Hanfling:2017:RPN

[HBD17] Dan Hanfling, Frederick M. Burkle, Jr., and Cham Dallas. The right planning now will save countless lives after a nuclear attack. Bulletin of the Atomic Scientists, 73(4):220–225, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1338005.

Hayes:2007:SD

[HBM^{+}07] Peter Hayes, Scott Bruce, Christina McCain, Norman Neureiter, and Eileen S. Vergino. Science as diplomacy. Bulletin of the Atomic Scientists, 63(4):22–23, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/22.

Hambleton:1971:ASM

[HC71] William W. Hambleton and Floyd L. Culler. Addenda [to salt mine disposal]. Bulletin of the Atomic Scientists, 27(6):31, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Huhn:1983:CWS

[HC83] Mary S. Huhn and Rosemary Chalk. Commentary: From women strike for peace. Bulletin of the Atomic Scientists, 39(2):51, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha82c] and comment [Dav83].

Hergen:1987:CSL

[HC87] Jim Hergen and Hodding Carter. Confidentially speaking … two views of the Reagan Administration’s via policy. Bulletin of the Atomic Scientists, 43(4):28, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hecker:2012:NKC

[HC12] Siegfried S. Hecker and Robert Carlin. North Korea in 2011: Countdown to Kim il-Sung’s centenary. Bulletin of the Atomic Scientists, 68(1):50–60, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/68/1/50.

Hoffman:2016:MCA

[HCH16] Robert R. Hoffman, Timothy M. Cullen, and John K. Hawley. The myths and costs of autonomous weapon systems. Bulletin of the Atomic Scientists, 72(4):247–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hackett:1982:FRB

[HD82] Ellen Hackett and John Dowling. Film review: The Race Nobody Wins, produced and directed by David Cartwright. Bulletin of the Atomic Scientists, 38(3):35, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heisbourg:1992:LMC

[HD92] François Heisbourg and Kim Dawson. Letters: More clock complaints. Bulletin of the Atomic Scientists, 48(4):47, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo91b].

He:2023:AAH

[HD23] Jingjie He and Nikita Degtyarev. AI and atoms: How artificial intelligence is revolutionizing nuclear material production. Bulletin of the Atomic Scientists, 79(5):316–328, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2245251.

Holdren:1972:ODE

[HE72] John P. Holdren and Paul R. Ehrlich. One-dimensional ecology revisited: A rejoinder. Bulletin of the Atomic Scientists, 28(6):

42–45, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [EH72a, Com72, EH72b].

Hutchins:1976:LMC

[HE76] Bruce A. Hutchins and John T. Edsall. Letter: Meaningless comparison. Bulletin of the Atomic Scientists, 32(10):8, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds76].

Head:1983:SIN

[Hea83] Ivan L. Head. Survival is the issue: North–South interdependence. Bulletin of the Atomic Scientists, 39(5):20–22, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hebert:1951:CB

[Heb51] F. Edward Hebert. The Congressman and the bomb. Bulletin of the Atomic Scientists, 7(7–8):219, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hebert:1952:LEA

[Heb52] F. Edward Hebert. Letter to the Editor: [atomic development]. Bulletin of the Atomic Scientists, 8(1):33, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hechter:1970:BBR

[Hec70] Oscar Hechter. Books: The Rise and Fall of T. D. Lysenko, by Zhores A. Medvedev. Bulletin of the Atomic Scientists, 26(4): 54–56, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heckman:1978:LSB

[Hec78] Thomas P. Heckman. Letter: Sexless breeders. Bulletin of the Atomic Scientists, 34(2):4–5, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heckrotte:1986:SVV

[Hec86] Warren Heckrotte. A Soviet view of verification. Bulletin of the Atomic Scientists, 42(8):12–15, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hecht:2006:SDT

[Hec06] Jeff Hecht. Space: Dual threat. Bulletin of the Atomic Scientists, 62(5):14–16, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/14.

Hecker:2008:DNK

[Hec08] Siegfried S. Hecker. Denuclearizing North Korea. Bulletin of the Atomic Scientists, 64(2):44–49, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/2/44.

Hecht:2012:EFU

[Hec12] Gabrielle Hecht. An elemental force: Uranium production in Africa, and what it means to be nuclear. Bulletin of the Atomic Scientists, 68(2):22–33, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/2/22.

Hecker:2016:QPC

[Hec16] Siegfried S. Hecker. Questions for the Presidential candidates on nuclear terrorism, proliferation, weapons policy, and energy. Bulletin of the Atomic Scientists, 72(5):276–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hecht:2019:SWS

[Hec19] Jeff Hecht. A “Star Wars” sequel? the allure of directed energy for space weapons. Bulletin of the Atomic Scientists, 75 (4):162–170, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1628507.

Hecht:2020:EFU

[Hec20] Gabrielle Hecht. 2012: An elemental force: Uranium production in Africa, and what it means to be nuclear. Bulletin of the Atomic Scientists, 76(6):431–437, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847526.

Hecht:2022:WWS

[Hec22] Jeff Hecht. Why will some promising renewables technologies enter a “Valley of Death,” from which they never emerge?

Bulletin of the Atomic Scientists, 79(1):3–8, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2155004.

Hedgpeth:1965:BHP

[Hed65] Joel W. Hedgpeth. Bodega Head — a partisan view. Bulletin of the Atomic Scientists, 21(3):2–7, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dav65] and reply [HH65].

Heisenberg:1947:RGT

[Hei47] Werner Heisenberg. Research in Germany on the technical application of atomic energy. Nature, 160(4059):211–215, August 16, 1947. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). See also [Hei53].

Heisenberg:1953:NP

[Hei53] Werner Heisenberg. Nuclear Physics. Philosophical Library, New York, NY, USA, 1953. ix + 225 pp. LCCN QC173 .H3854. See also [Hei47].

Heiberg:1954:BBP

[Hei54] Edvard Heiberg. Books: Nuclear Physics, by Werner Heisenberg. Bulletin of the Atomic Scientists, 10(6):233, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heitler:1964:FES

[Hei64] W. Heitler. Forum: Ethics of the scientific age. Bulletin of the Atomic Scientists, 20(8):21–23, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heisenberg:1968:TRA

[Hei68] Werner Heisenberg. The Third Reich and the atomic bomb. Bulletin of the Atomic Scientists, 24(6):34–35, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://alsos.wlu.edu/information.aspx?id=

157. Translated by Margaret Seckel from the German original.

Heisenberg:1970:GAA

[Hei70] Werner Heisenberg. Genesis of the atomic age: Physics and beyond: Encounters and conversations: The responsibility of the scientist (1945–1950). Bulletin of the Atomic Scientists, 26(9):33– 36, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from [Hei71].

Heisenberg:1971:PBE

[Hei71] Werner Heisenberg. Physics and beyond; encounters and conversations, volume 42 of World perspectives. Harper & Row, New York, NY, USA, 1971. xviii + 247 pp. LCCN QC173 .H38613 1971. Translated from the German by Arnold J. Pomerans.

Heisenberg:1973:TSS

[Hei73] Werner Heisenberg. Tradition in science. Science and public affairs. Bulletin of the Atomic Scientists, 29(10):4–10, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heinrich:1989:LTU

[Hei89] Dieter Heinrich. Letter: Trust U.N. to verify. Bulletin of the Atomic Scientists, 45(4):53–54, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol88e] and comment [Sco89].

Heinonen:2011:MFC

[Hei11] Olli Heinonen. A multinational fuel consortium: Obstacles, options, and ways forward. Bulletin of the Atomic Scientists, 67 (4):37–42, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/4/37.

Hekhuis:1960:GPP

[Hek60] Dale J. Hekhuis. Goals, plans, and priorities in Soviet–American cooperation. Bulletin of the Atomic Scientists, 16(3):97–102, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Helliwell:1970:UAS

[Hel70] Robert A. Helliwell. The upper atmosphere as seen from Antarctica. Bulletin of the Atomic Scientists, 26(10):55–61, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Helmer:1977:LAH

[Hel77a] Borden Helmer. Letter: Arms and human rights. Bulletin of the Atomic Scientists, 33(7):2, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Yoa78].

Helmer:1977:LMT

[Hel77b] Borden Helmer. Letter: Mankind as trustee. Bulletin of the Atomic Scientists, 33(9):2–3, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar77c].

Hellerstein:1985:LPN

[Hel85] Nathaniel S. Hellerstein. Letter: Planer or nation? Bulletin of the Atomic Scientists, 41(2):54, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gor84].

Helton:2002:USA

[Hel02] Arthur C. Helton. Unpleasant surprises await. Bulletin of the Atomic Scientists, 58(6):94–100, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/6/94.

Heller:2007:OIO

[Hel07] Steven Heller. The other Icon of our age. Bulletin of the Atomic Scientists, 63(1):22–23, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/1/22.

Hellman:2011:HRN

[Hel11] Martin E. Hellman. How risky is nuclear optimism? Bulletin of the Atomic Scientists, 67(2):47–56, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/67/2/47.

Henkin:1959:AIC

[Hen59] Louis Henkin. Arms inspection and the Constitution. Bulletin of the Atomic Scientists, 15(5):192–197, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Henoch:1968:CT

[Hen68] Ruth L. Henoch. The chemistry of time. Bulletin of the Atomic Scientists, 24(10):38–41, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Henoch:1971:BBS

[Hen71a] Ruth Henoch. Books: The Second Genesis: The Coming Control of Life, by Albert Rosenfeld. Bulletin of the Atomic Scientists, 27(1):47–48, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Henry:1971:CQ

[Hen71b] Norman Henry. Communications: A question. Bulletin of the Atomic Scientists, 27(1):52, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Henderson:1978:CTF

[Hen78a] Carter Henderson. Commentary: The tragic failure of energy planning. Bulletin of the Atomic Scientists, 34(10):15–19, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Henderson:1978:FP

[Hen78b] Carter Henderson. The frugality phenomenon. Bulletin of the Atomic Scientists, 34(5):24–27, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Henry:1984:CMR

[Hen84] Norman Henry. Commentary: Military rituals. Bulletin of the Atomic Scientists, 40(8):44, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [PG84].

Henderson:1993:WCD

[Hen93] Simon Henderson. “We can do it ourselves”. Bulletin of the Atomic Scientists, 49(7):27–32, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

HLSC:1996:FSP

[Hen96] Henry L.Stimson Center. A four-step program to nuclear disarmament. Bulletin of the Atomic Scientists, 52(2):52–55, March/ April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heneghan:2017:PBP

[Hen17] Liam Heneghan. Preserving biodiversity, preventing climate disaster: Childish dreams or audacious strategies? Bulletin of the Atomic Scientists, 73(4):284–287, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1338050.

Heppe:1977:LMB

[Hep77] Robert Heppe. Letter: Mind-boggling. Bulletin of the Atomic Scientists, 33(5):2, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herter:1960:MKC

[Her60] Christian A. Herter. [Mr. Khrushchev, a common language, and a common interest]. Bulletin of the Atomic Scientists, 16(1):13, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Three-paragraph excerpt from a speech before the National Foreign Trade Council, 16 November 1959.

Herskowitz:1967:FWG

[Her67] D. S. Herskowitz. Forum: Wheat gap? Bulletin of the Atomic Scientists, 23(6):64, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab67c].

Hermann:1978:LES

[Her78a] Cal C. Hermann. Letter: Energy from salt domes. Bulletin of the Atomic Scientists, 34(6):5, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hersfeld:1978:CMR

[Her78b] Charles M. Hersfeld. Commentary: The military R&D process — a view from industry. Bulletin of the Atomic Scientists, 34 (10):33–40, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herman:1979:CTD

[Her79] Theodore Herman. Commentary: Teaching disarmament and alternative security systems. Bulletin of the Atomic Scientists, 35(4):54, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1982:BWW

[Her82a] Gregg Herken. From The Winning Weapon: Epilogue. Bulletin of the Atomic Scientists, 38(10):36–37, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hertsgaard:1982:BRB

[Her82b] Mark Hertsgaard. Book review: The Nuclear Barons, by Peter Pringle and James Spigelman. Bulletin of the Atomic Scientists, 38(4):41–42, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1983:BRB

[Her83a] Gregg Herken. Book reviews: Power and Principle: Memoirs of the National Security Adviser, 1977–1981, by Zbigniew Brzezinski; Hard Choices: Four Critical Years in American Foreign Policy, by Cyrus Vance. Bulletin of the Atomic Scientists, 39(10):39– 40, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herzig:1983:CRR

[Her83b] Christopher Herzig. Commentary: Rajasthan reactor. Bulletin of the Atomic Scientists, 39(3):64, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil82c].

Hertsgaard:1984:SNP

[Her84] Mark Hertsgaard. Should nuclear power be saved? Bulletin of the Atomic Scientists, 40(8):30–31, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1985:BRB

[Her85] Gregg Herken. Book review: Strategic Stalemate: Nuclear Weapons and Arms Control in American Politics, by Michael Krepon. Bulletin of the Atomic Scientists, 41(10):51–52, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1987:EOS

[Her87] Gregg Herken. The earthly origins of Star Wars. Bulletin of the Atomic Scientists, 43(8):20–28, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herby:1988:PUD

[Her88a] Peter Herby. Perspective: U.N. disarmament session fizzles. Bulletin of the Atomic Scientists, 44(7):6–7, September 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1988:BRB

[Her88b] Gregg Herken. Book review: Science and Technology Advice to the President, Congress, and Judiciary , edited by William T. Golden. Bulletin of the Atomic Scientists, 44(7):56–57, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herolf:1988:NTF

[Her88c] Gunilla Herolf. New technology favors defense. Bulletin of the Atomic Scientists, 44(7):42–44, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1990:BRN

[Her90a] Gregg Herken. Book review: Nuclear zookeeper: Monkeys, Men, and Missiles: An Autobiography, 1946–1988, by Solly Zuckerman. Bulletin of the Atomic Scientists, 46(3):50–51, April 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:1990:BRT

[Her90b] Gregg Herken. Book review: Teller retells: Edward Teller: Giant of the Golden Age of Physics, by Stanley A. Blumberg and Louis G. Panos. Bulletin of the Atomic Scientists, 46(8):40–41, October 1990. CODEN BASIAP. ISSN 0096-3402.

Herby:1995:GOO

[Her95a] Peter Herby. Guest opinion: Outlaw blinding. Bulletin of the Atomic Scientists, 51(2):4, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herman:1995:LNN

[Her95b] Edward S. Herman. Letter: Neither new nor benign. Bulletin of the Atomic Scientists, 51(5):59, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo95j].

Hershberg:2002:BO

[Her02] James G. Hershberg. A brilliant organizer. Bulletin of the Atomic Scientists, 58(3):63–65, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/3/63.

Hester:1955:LL

[Hes55] A. G. Hester. Letter from London. Bulletin of the Atomic Scientists, 11(5):187–190, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hester:1978:ICS

[Hes78] James M. Hester. An international community of scholars: the newly-formed United Nations University is mobilizing worldwide science and scholarship to help solve worldwide problems. Bulletin of the Atomic Scientists, 34(3):36–41, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hesburgh:1984:HEN

[Hes84] Theodore M. Hesburgh. Higher education and the nuclear crisis. Bulletin of the Atomic Scientists, 40(10):3s–3s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hesburgh:1985:RSL

[Hes85] Theodore Hesburgh. Religious, scientific leaders on arms race. Bulletin of the Atomic Scientists, 41(4):49, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hesketh:1997:LCN

[Hes97] R. V. Hesketh. Letter: Chernobyl’s nuclear secret. Bulletin of the Atomic Scientists, 53(2):61, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sic96a] and comment [Sta97].

Hetherington:1955:BVA

[Het55] Alastair Hetherington. A British view of American security policy. Bulletin of the Atomic Scientists, 11(4):161–163, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Hevesy:1947:LBE

[Hev47] G. Hevesy. Lifting the ban on export of tracers. Bulletin of the Atomic Scientists, 3(11):320, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hewlett:1981:BCA

[Hew81] Richard G. Hewlett. ‘Born classified’ in the AEC: A historian’s view. Bulletin of the Atomic Scientists, 37(10):20–27, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Heyerdahl:1972:VR

[Hey72] Thor Heyerdahl. A view from a raft. Bulletin of the Atomic Scientists, 28(7):52–53, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hogan:1980:CEG

[HF80] James P. Hogan and Bernard T. Feld. Commentary: Enough gloom and doom. Bulletin of the Atomic Scientists, 36(8):63– 64, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hallenbeck:1989:LWA

[HFMH89] Ralph A. Hallenbeck, Joseph Forbes, Jack Mendelsohn, and Thomas Halverson. Letters: Who’s ahead in Europe? Bulletin of the Atomic Scientists, 45(6):52–54, July/August 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [MH89a].

Herman:1979:HRU

[HFMS79] Edward S. Herman, David Frisch, Robert Michaels, and Daniel B. Schirmer. On human rights and university contracts. Bulletin of the Atomic Scientists, 35(7):60–63, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fri79].

Hankins:1963:LEA

[HFO63] Frank H. Hankins, Irving H. Flamm, and Warren E. Olson. Letters to the Editor: The approach to responsibility. Bulletin of the Atomic Scientists, 19(6):29–30, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ols63].

Horst:1971:CLC

[HG71] David R. Horst and Bernard Greenblatt. Communications: Licensing: For cars and babies: Comment on Russett. Bulletin of the Atomic Scientists, 27(5):2–3, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rus70a, Rus71].

Hamburg:1984:NCM

[HG84] David A. Hamburg and Alexander L. George. Nuclear crisis management. Bulletin of the Atomic Scientists, 40(6):24–28, June/ July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harrington:2017:WAM

[HGF17] Anne I. Harrington, Eliza Gheorghe, and Anya Loukianova Fink. What arguments motivate citizens to demand nuclear disarmament? Bulletin of the Atomic Scientists, 73(4):255–263, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1338039.

Horovitz:2010:BCP

[HGV10] Liviu Horovitz and Robert Golan-Vilella. Boosting the CTBT’s prospects in the Middle East. Bulletin of the Atomic Scientists, 66(2):9–16, March 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/2/9.

Hedgpeth:1965:LEB

[HH65] Joel W. Hedgpeth and Garrett Hardin. Letter to the Editor: Bodega Head. Bulletin of the Atomic Scientists, 21(10):23, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav65].

Halasz:1975:TT

[HH75] Nicholas Halasz and Robert Halasz. The tables are turned. Bulletin of the Atomic Scientists, 31(1):36, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harwell:1987:UNW

[HH87] Mark A. Harwell and Christine C. Harwell. Updating the “nuclear winter” debate. Bulletin of the Atomic Scientists, 43(8):42–44, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:2007:SA

[HHH07] Chuck Hansen, Eleanor Hansen, and Larry Hatfield, editors. The Swords of Armageddon. Chukelea Publications, Sunnyvale, CA, USA, second edition, 2007. ISBN 0-9791915-0-5 (CD-ROM). various pp. LCCN U264.3 .S86 2007. One CD-ROM.

Hoffman:1972:BBL

[HI72] Allan R. Hoffman and David Rittenhouse Inglis. Books: Low-Level Radiation, by Ernest J. Sternglass. Bulletin of the Atomic Scientists, 28(10):45–52, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herman:1994:LBW

[HI94] Edward S. Herman and John Isaacs. Letter: Bogged down in the wrong quagmire? Bulletin of the Atomic Scientists, 50(6):3, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa94d].

Hellman:2005:CT

[HI05a] Christopher Hellman and John Isaacs. Closing time. Bulletin of the Atomic Scientists, 61(6):20–75, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hellman:2005:MCT

[HI05b] Christopher Hellman and John Isaacs. Military: Closing time. Bulletin of the Atomic Scientists, 61(6):20–21, November/ December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
6/20.

Hiatt:1980:PFE

[Hia80] Howard H. Hiatt. Preventing the final epidemic. Bulletin of the Atomic Scientists, 36(7):15–16, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hiatt:1987:PHC

[Hia87] Howard H. Hiatt. Perspective: Health care bites the bullet. Bulletin of the Atomic Scientists, 43(5):7–8, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1968:CCA

[Hib68] A. R. Hibbs. Current comments: ABM and the algebra of uncertainty. Bulletin of the Atomic Scientists, 24(3):31–33, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bre68].

Hibbs:1994:A

[Hib94a] Mark Hibbs. “Agent 008”. Bulletin of the Atomic Scientists, 50(6):28, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1994:BPP

[Hib94b] Mark Hibbs. Bulletin: Plutonium powder puzzles police. Bulletin of the Atomic Scientists, 50(5):6–7, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1994:PF

[Hib94c] Mark Hibbs. Plutonium fingerprints. Bulletin of the Atomic Scientists, 50(6):27, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1994:PPP

[Hib94d] Mark Hibbs. Plutonium, politics and panic. Bulletin of the Atomic Scientists, 50(6):24–31, November/December 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1994:SA

[Hib94e] Mark Hibbs. A second agenda. Bulletin of the Atomic Scientists, 50(6):30, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1995:BAB

[Hib95a] Mark Hibbs. Bulletin: Attali bombshell a real dud. Bulletin of the Atomic Scientists, 51(5):5–6, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1995:BFT

[Hib95b] Mark Hibbs. Bulletin: “Fairy tales” in Munich. Bulletin of the Atomic Scientists, 51(6):5–7, November/December 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1995:BQG

[Hib95c] Mark Hibbs. Bulletin: Questioning German intelligence. Bulletin of the Atomic Scientists, 51(4):5–6, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1995:BSS

[Hib95d] Mark Hibbs. Bulletin: Siemens steps into the breach. Bulletin of the Atomic Scientists, 51(1):6–7, January/February 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SH95].

Hibbs:1995:RNSa

[Hib95e] Mark Hibbs. Report: Nuclear smuggling: Czechs seize migrating uranium. Bulletin of the Atomic Scientists, 51(2):13–14, March/ April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1995:RNSb

[Hib95f] Mark Hibbs. Report: Nuclear smuggling: Which fissile fingerprint? Bulletin of the Atomic Scientists, 51(3):10–11, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1996:BGK

[Hib96a] Mark Hibbs. Bulletin: Germans knew plutonium sting was “problematic”. Bulletin of the Atomic Scientists, 52(2):7–8, March/ April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hibbs:1996:TE

[Hib96b]  Mark Hibbs. Tomorrow, a Eurobomb?  Bulletin of the

Atomic Scientists, 52(1):16–23, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [FC96, Kün96].

Hibbs:2001:PCG

[Hib01] Mark Hibbs. Plutonium: Can Germany swear off? Bulletin of the Atomic Scientists, 57(3):63–66, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/63.

Hibbs:2006:UNS

[Hib06] Mark Hibbs. The unmaking of a nuclear smuggler. Bulletin of the Atomic Scientists, 62(6):35–41, November/December 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/6/35.

Hibbs:2012:NEW

[Hib12] Mark Hibbs. Nuclear energy 2011: a watershed year. Bulletin of the Atomic Scientists, 68(1):10–19, January/February 2012. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/68/1/10.

Hickey:2022:TVH

[Hic22] Samuel M. Hickey. Trust but verify: How to get there by using next-generation nuclear verification and warhead dismantlement techniques. Bulletin of the Atomic Scientists, 78(2):91– 97, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2038895.

Higinbotham:1946:RS

[Hig46] William Higinbotham. The road to security. Bulletin of the Atomic Scientists, 2(3–4):27–28, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higginbotham:1948:MVC

[Hig48a] William A. Higginbotham. Military vs. civilian control of bomb debated in secret. Bulletin of the Atomic Scientists, 4(9):260, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higinbotham:1948:FAS

[Hig48b] W. A. Higinbotham. The Federation of American Scientists. Bulletin of the Atomic Scientists, 4(1):21–22, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higinbotham:1950:SDW

[Hig50] William A. Higinbotham. Scientists discuss war and peace. Bulletin of the Atomic Scientists, 6(11):350, November 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higinbotham:1963:LEA

[Hig63] W. A. Higinbotham. Letter to the Editor: AEC history. Bulletin of the Atomic Scientists, 19(2):34–35, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higinbotham:1966:BBP

[Hig66] William A. Higinbotham. Books: A Peril and a Hope (The Scientists’ Movement in America, 1945–47), by Alice Kimball Smith; The Decision to Drop the Bomb (a Political History), by Len Giovannitti and Fred Freed; Atomic Diplomacy: Hiroshima and Potsdam, by Gar Alperovitz. Bulletin of the Atomic Scientists, 22 (4):34–37, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higgins:1976:LJP

[Hig76] John L. B. Higgins. Letter: JFK’s profound grasp. Bulletin of the Atomic Scientists, 32(5):2–3, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber76b].

Highland:1977:FSC

[Hig77] Virgil L. Highland. The fire safety controversy. Bulletin of the Atomic Scientists, 33(8):54–55, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rob78a].

Higinbotham:1989:LIM

[Hig89] William A. Higinbotham. Letter: IAEA maligned …. Bulletin of the Atomic Scientists, 45(2):46, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hjalmarsson:2010:GWS

[HIKWB10] Karin Hjalmarsson, Nicolas Isla, Gabriele Kraatz-Wadsack, and Maurizio Barbeschi. Global watch: The state of biological investigations. Bulletin of the Atomic Scientists, 66(4):70–76, July/ August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
4/70.

Hill:1946:MRS

[Hil46] A. V. Hill. The moral responsibilities of scientists. Bulletin of the Atomic Scientists, 1(7):3, 15, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hill:1951:SQO

[Hil51] A. V. Hill. I. Scientists are quite ordinary folk. Bulletin of the Atomic Scientists, 7(12):371–372, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hill:1952:EDS

[Hil52] A. V. Hill. The ethical dilemma of science. Bulletin of the Atomic Scientists, 8(8):262–266, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hildebrand:1953:PHP

[Hil53] Joel H. Hildebrand. The professor and his public. Bulletin of the Atomic Scientists, 9(1):23–25, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hilsman:1960:PNS

[Hil60] Roger Hilsman. Planning for national security: A proposal. Bulletin of the Atomic Scientists, 16(3):93–96, 112, March 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hill:1965:ICP

[Hil65a] Gladwin Hill. An industry combats pollution. Bulletin of the Atomic Scientists, 21(10):42–43, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hilsman:1965:PPS

[Hil65b] Roger Hilsman. The policy proposals: “Strength and conciliation”. Bulletin of the Atomic Scientists, 21(6):41–42, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hill:2012:LDP

[Hil12] Colin K. Hill. The low-dose phenomenon: How bystander effects, genomic instability, and adaptive responses could transform cancer-risk models. Bulletin of the Atomic Scientists, 68 (3):51–58, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/3/51.

Hill:2016:POM

[Hil16] Fiona Hill. Putin: The one-man show the West doesn’t understand. Bulletin of the Atomic Scientists, 72(3):140–144, 2016.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hill:2020:POM

[Hil20] Fiona Hill. 2016: Putin: The one-man show the West doesn’t understand. Bulletin of the Atomic Scientists, 76(6):438–442, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847532.

Hinton:1953:AED

[Hin53] Sir Christopher Hinton. Atomic energy developments in Great Britain. Bulletin of the Atomic Scientists, 9(10):366–368, 390, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hiner:1960:LEH

[Hin60] James Hiner. Letter to the Editor: “High noon”? Bulletin of the Atomic Scientists, 16(6):257, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hinshelwood:1967:SS

[Hin67] Cyril Hinshelwood. Science and scientists. Bulletin of the Atomic Scientists, 23(6):30–36, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hines:1969:WSN

[Hin69a] William Hines. Washington scene: The news under new management? Bulletin of the Atomic Scientists, 25(3):38, 42, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hines:1969:WSU

[Hin69b] William Hines. Washington scene: Unhappy landings: The mess in civil aviation. Bulletin of the Atomic Scientists, 25(5):37–40, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hinderstein:2018:IPN

[Hin18] Corey Hinderstein. International partnership for nuclear disarmament verification: Laying a foundation for future arms reductions. Bulletin of the Atomic Scientists, 74(5):305–311, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1507785.

Hirschfelder:1950:EAW

[Hir50] Joseph O. Hirschfelder. The effects of atomic weapons. Bulletin of the Atomic Scientists, 6(8–9):236–240, 285–286, August/ September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Hirschfelder:1955:OSM

[Hir55] Joseph A. Hirschfelder. Organization scientists to meet a national emergency. Bulletin of the Atomic Scientists, 11(6):223–224, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1959:GE

[Hir59] Étienne Hirsch. A guide to Euratom. Bulletin of the Atomic Scientists, 15(6):250–252, 265, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from The New Scientist.

Hirsch:1965:KW

[Hir65] Walter Hirsch. Knowledge for what? Bulletin of the Atomic Scientists, 21(5):28–30, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1972:CUF

[Hir72] Abraham M. Hirsch. Can the university forget the past and find happiness in the present? Bulletin of the Atomic Scientists, 28 (9):11–15, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirst:1973:TEU

[Hir73] Eric Hirst. Transportation energy use and conservation potential. Bulletin of the Atomic Scientists, 29(9):36–42, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirschberg:1976:PPS

[Hir76] Alan S. Hirschberg. Public policy for solar heating and cooling. Bulletin of the Atomic Scientists, 32(8):37–45, October 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rom77].

Hirst:1981:CCC

[Hir81] Eric Hirst. Commentary: Cutting into conservation. Bulletin of the Atomic Scientists, 37(5):10, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1988:BRB

[Hir88] Daniel Hirsch. Book review: Better a Shield than a Sword: Perspectives on Defense and Technology, by Edward Teller. Bulletin of the Atomic Scientists, 44(4):48–49, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1989:LLE

[Hir89a] Helmut Hirsch. Letter: Long enough to make a bomb. Bulletin of the Atomic Scientists, 45(9):50, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fis89].

Hirsch:1989:LWI

[Hir89b] Helmut Hirsch. Letter: Or was it? Bulletin of the Atomic Scientists, 45(2):46–47, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bol88a] and comment [Fis89].

Hirsch:2002:NWM

[Hir02] Daniel Hirsch. The NRC: What me worry? Bulletin of the Atomic Scientists, 58(1):38–44, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/38.

Hitchens:2001:GPP

[Hit01] Theresa Hitchens. Get a policy, please. Bulletin of the Atomic Scientists, 57(1):21–23, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/1/21.

Hitchens:2003:ACE

[Hit03] Theresa Hitchens. Article commentary: Everyone will want one. Bulletin of the Atomic Scientists, 59(1):22–23, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/1/22.

Hitchens:2007:RSW

[Hit07] Theresa Hitchens. Return of the star warriors. Bulletin of the Atomic Scientists, 63(1):76–78, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/1/76.

Hejl:1976:LBC

[HJ76] Lubos G. Hejl and František Janouch. Letter: Boomerang in Czechoslovakia. Bulletin of the Atomic Scientists, 32(9):3, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jan76c].

Horshead:1992:LRC

[HJS^{+}92] Mike Horshead, Courtney Jones, Arnold Simoni, John M. Lamb, Bradford Lyttle, and Victor Reilly. Letter: Readers call clock move cuckoo. Bulletin of the Atomic Scientists, 48(2):45–47, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo91b].

Hoodbhoy:1996:NST

[HK96] Pervez Hoodbhoy and Martin Kalinowski. The nuclear subcontinent: The tritium solution. Bulletin of the Atomic Scientists, 52 (4):41–44, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Higham:2002:MHS

[HK02] Robin D. S. Higham and Frederick W. Kagan, editors. The military history of the Soviet Union. Palgrave, New York, NY, USA, 2002. ISBN 0-312-29398-4. vii + 328 pp. LCCN DK54 .M49 2002.

URL http://www.loc.gov/catdir/bios/hol058/2002265813.
html;   http://www.loc.gov/catdir/description/hol052/
2002265813.html;   http://www.loc.gov/catdir/toc/fy02/
2002265813.html.

Holt:2009:FPF

[HK09] Lynne Holt and Theodore J. Kury. Florida’s plans to finance new nuclear plants. Bulletin of the Atomic Scientists, 65(4):31–40, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/4/31.

Hultman:2013:TMI

[HK13] Nathan Hultman and Jonathan Koomey. Three Mile Island: The driver of US nuclear power’s decline? Bulletin of the Atomic Scientists, 69(3):63–70, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/3/63.

Hill:2018:ORC

[HK18] Alice C. Hill and William Kakenmaster. An overview of “resilience” and climate change. Bulletin of the Atomic Scientists, 74(2):61–65, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2018.1436803.

Horowitz:2021:HJB

[HK21] Michael C. Horowitz and Lauren Kahn. How Joe Biden can use confidence-building measures for military uses of AI. Bulletin of the Atomic Scientists, 77(1):33–35, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1860331.

Haase:1978:LNA

[HKC78] Beth Haase, Chris Kettering, and Michelle Carrade. Letter: A necessary alternative. Bulletin of the Atomic Scientists, 34(4):4, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Harrison:1949:SAI

[HKL^{+}49] George R. Harrison, Paul E. Klopsteg, F. W. Loomis, George B. Pegram, and Wallace Waterfall. Statement by the American Institute of Physics. Bulletin of the Atomic Scientists, 5(6–7):176, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hickenlooper:1949:HVL

[HL49] Bourke B. Hickenlooper and David E. Lilienthal. Hickenlooper versus Lilienthal. Bulletin of the Atomic Scientists, 5(6–7):182– 183, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hlavka:1965:LES

[Hla65] George E. Hlavka. Letter to the Editor: Siblings of the test ban. Bulletin of the Atomic Scientists, 21(9):24, November 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:2003:NDL

[HLL03] Daniel Hirsch, David Lochbaum, and Edwin Lyman. The NRC’s dirty little secret. Bulletin of the Atomic Scientists, 59(3):44–51, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
3/44.

Humphrey:1972:MIM

[HM72] Hubert H. Humphrey and George McGovern. Men and issues: McGovern and Humphrey discuss science policy, environment, energy and space. Bulletin of the Atomic Scientists, 28(6):46–49, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1990:HBW

[HM90a] Daniel Hirsch and William G. Mathews. The H-bomb: Who really gave away the secret? Bulletin of the Atomic Scientists, 46(1):22–30, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kno90, MHM90, BHM91, HM97].

Hirsch:1990:WWK

[HM90b] Daniel Hirsch and William G. Mathews. Who was Klaus Fuchs? Bulletin of the Atomic Scientists, 46(1):26, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1997:LHB

[HM97] Daniel Hirsch and William G. Mathews. Letter: H-bomb secrets. Bulletin of the Atomic Scientists, 53(4):3, 58, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90a, KAS96] and comment [Han97].

Hoodbhoy:1998:SLE

[HM98] Pervez Hoodbhoy and Zia Mann. Sanctions: Lift ’em. Bulletin of the Atomic Scientists, 54(5):20–21, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hecker:2015:EAO

[HM15] Siegfried S. Hecker and Abbas Milani. Ending the assassination and oppression of Iranian nuclear scientists. Bulletin of the Atomic Scientists, 71(1):46–52, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hirsch:1986:PRT

[HMR86] Daniel Hirsch, Stephanie Murphy, and Bennett Ramberg. Protecting reactors from terrorists. Bulletin of the Atomic Scientists, 42(3):22–25, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hurwitz:1981:CIA

[HNB81] Henry Hurwitz, Jr., Anthony V. Nero, Jr., and Jan Beyea. Commentary: Indoor air pollution. Bulletin of the Atomic Scientists, 37(2):61–64, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hinderstein:2012:HME

[HNR12] Corey Hinderstein, Andrew Newman, and Ole Reistad. From HEU minimization to elimination: Time to change the vocabulary. Bulletin of the Atomic Scientists, 68(4):83–95, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/4/83.

Halperin:1988:NF

[HO88] Morton H. Halperin and Madalene O’Donnell. The nuclear fallacy. Bulletin of the Atomic Scientists, 44(1):6–11, January/ February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Loe88].

Hoag:1957:DPM

[Hoa57] Malcolm W. Hoag. Is “Dual” preparedness more expensive? Bulletin of the Atomic Scientists, 13(2):48–51, February 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoagland:1959:SRS

[Hoa59] Hudson Hoagland. Some reflections on science and society. Bulletin of the Atomic Scientists, 15(7):284–287, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoagland:1964:CPC

[Hoa64] Hudson Hoagland. Cybernetics of population control. Bulletin of the Atomic Scientists, 20(2):2–6, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoagland:1969:TAE

[Hoa69] Hudson Hoagland. Technology, adaptation, and evolution. Bulletin of the Atomic Scientists, 25(1):27–30, January 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoagland:1972:CMC

[Hoa72] Hudson Hoagland. Can man control his biological evolution? — a symposium on genetic engineering: I. Man’s responsibility to his future. Bulletin of the Atomic Scientists, 28(10):13, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoagland:1977:BEB

[Hoa77] Hudson Hoagland. Brain evolution and the biology of belief: a scientist reflects on man’s drive toward self-destruction. Bulletin of the Atomic Scientists, 33(3):41–44, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Har77a].

Hobson:1978:LSL

[Hob78] Art Hobson. Letter: Sexist language. Bulletin of the Atomic Scientists, 34(6):5, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hobson:1985:BRB

[Hob85] Art Hobson. Book review: Missiles for the Nineties: ICBMs and Strategic Policy, edited by Barry R. Schneider, Colin S. Gray, and Keith B. Payne. Bulletin of the Atomic Scientists, 41(9): 46–48, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hocking:1957:LES

[Hoc57] William Ernest Hocking. Letter to the Editor: Soviet scientists’ reaction to the Canada Conference. Bulletin of the Atomic Scientists, 13(10):377, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab57c, Kra57].

Hodgson:1955:ICA

[Hod55] P. E. Hodgson. International control of atomic energy. Bulletin of the Atomic Scientists, 11(2):63, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hodgson:1959:BAS

[Hod59] P. E. Hodgson. The British Atomic Scientists’ Association, 1946– 59. Bulletin of the Atomic Scientists, 15(9):392–394, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hodge:1969:BD

[Hod69] Carle O. Hodge. The blooming desert. Bulletin of the Atomic Scientists, 25(9):32–33, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hodgkin:1981:U

[Hod81a] Dorothy Hodgkin. It’s up to us! Bulletin of the Atomic Scientists, 37(1):38–39, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See errata [Ano81e] and comment [Hod81b].

Hodgkin:1981:CLH

[Hod81b] Dorothy C. Hodgkin. Commentary: Lessons from history. Bulletin of the Atomic Scientists, 37(9):56–57, November 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hod81a].

Hodgson:2000:CLW

[Hod00] Martin Hodgson. The coca leaf war. Bulletin of the Atomic Scientists, 56(3):36–45, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/3/36.

Hoenig:1987:EDB

[Hoe87] Milton M. Hoenig. Energy Department blurs line between civilian, military reactors. Bulletin of the Atomic Scientists, 43(5): 25–27, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bea87].

Hoekema:1988:BRB

[Hoe88] Jan Th. Hoekema. Book review: The International Atomic Energy Agency and World Nuclear Order, by Lawrence Scheinman. Bulletin of the Atomic Scientists, 44(10):45–46, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoey:2009:IQD

[Hoe09] Matthew Hoey. India’s quest for dual-use technology. Bulletin of the Atomic Scientists, 65(5):43–59, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/5/43.

Hoffman:1961:BS

[Hof61] Paul G. Hoffman. Bootstrap statemanship. Bulletin of the Atomic Scientists, 17(10):412–414, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoffman:1986:CES

[Hof86] Erik P. Hoffman. Chernobyl: the emerging story: Nuclear deception: Soviet information policy. Bulletin of the Atomic Scientists, 42(7):32–37, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoffmann:1988:GTR

[Hof88] Erik P. Hoffmann. Gorbachev’s trade reforms: something old, something new. Bulletin of the Atomic Scientists, 44(5):22–25, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoge:1968:BBA

[Hog68] James Hoge. Books: The Artillery of the Press, by James Reston. Bulletin of the Atomic Scientists, 24(1):53–54, January 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hogan:1986:LSB

[Hog86] Bill Hogan. Letter: Stopping the B-1. Bulletin of the Atomic Scientists, 42(3):57, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pai85e].

Hogendoorn:1997:CWA

[Hog97] E. J. Hogendoorn. A chemical weapons atlas. Bulletin of the Atomic Scientists, 53(5):35–39, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hohenemser:1966:ST

[Hoh66] Kurt H. Hohenemser. The supersonic transport. Bulletin of the Atomic Scientists, 22(10):8–12, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hohenemser:1975:LPC

[Hoh75] Kurt H. Hohenemser. Letter: Passive civil defense. Bulletin of the Atomic Scientists, 31(9):4, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hollitscher:1953:DMP

[Hol53] Walter Hollitscher. Dialectical materialism and the physicist. Bulletin of the Atomic Scientists, 9(2):54–57, March 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Physikalische Blätter/.

Hollaender:1956:MRR

[Hol56] Alexander Hollaender. Modification of radiation response. Bulletin of the Atomic Scientists, 12(3):76–80, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Ano56a]. This is a condensed version of a longer paper presented at the International Conference on the Peaceful Uses of Atomic Energy in Geneva on 17 August 1955.

Holifield:1958:RMB

[Hol58] Chet Holifield. Radiation and man: Broad studies: Congressional hearings on radioactive fall-out. Bulletin of the Atomic Scientists, 14(1):52–54, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holty:1959:AC

[Hol59] Carl Holty. Artistic creativity. Bulletin of the Atomic Scientists, 15(2):77–81, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holton:1962:PWB

[Hol62] Gerald Holton. Percy Williams Bridgman. Bulletin of the Atomic Scientists, 18(2):22–23, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holman:1964:PRD

[Hol64] Mary A. Holman. Patents for R and D. Bulletin of the Atomic Scientists, 20(8):32–34, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holland:1965:KU

[Hol65] Kenneth Holland. For knowledge and understanding. Bulletin of the Atomic Scientists, 21(4):45–46, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holmes:1966:MDN

[Hol66] Robert L. Holmes. Moral decision in the Nuclear Age. Bulletin of the Atomic Scientists, 22(4):27–29, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1973:CLR

[Hol73] John P. Holdren. Communications: Long-range forecasts. Bulletin of the Atomic Scientists, 29(9):2, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi73].

Holdren:1974:HNF

[Hol74] John P. Holdren. Hazards of the nuclear fuel cycle: The solutions to the problem lie beyond technology. Bulletin of the Atomic Scientists, 30(8):14–23, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Coh75a].

Holdren:1975:EPS

[Hol75] John P. Holdren. Energy and prosperity: Some elements of a global perspective. Bulletin of the Atomic Scientists, 31(1):25– 28, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mus75].

Holdren:1976:NCL

[Hol76a] John P. Holdren. The nuclear controversy and limitations of decision-making by experts. Bulletin of the Atomic Scientists, 32(3):20–22, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Coo76c].

Holton:1976:RMT

[Hol76b] Gerald Holton. Reflections on modern terrorism. Bulletin of the Atomic Scientists, 32(9):8–9, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Holdren:1977:BSB

[Hol77] John P. Holdren. Books: A strategy to bury time: New policy study provides some intellectual underpinning for Carter’s energy program: Nuclear Power: Issues and Choices. Bulletin of the Atomic Scientists, 33(6):58–63, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mac77c].

Holdren:1978:FPN

[Hol78] J. P. Holdren. Fusion power and nuclear weapons: a significant link? Bulletin of the Atomic Scientists, 34(3):4–5, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Hur78b, Gom78].

Holdren:1980:ELN

[Hol80a] John P. Holdren. Environmental liabilities of nuclear power. Bulletin of the Atomic Scientists, 36(1):31–32, January 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holzman:1980:DRC

[Hol80b] Franklyn D. Holzman. Dollars or rubles: the CIA’s military estimates. Bulletin of the Atomic Scientists, 36(6):23–27, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hollins:1982:DWS

[Hol82] Harry B. Hollins. A defensive weapons system. Bulletin of the Atomic Scientists, 38(6):63–65, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1983:BRR

[Hol83a] John P. Holdren. Book review: The risk assessors: Risk/Benefit Analysis, by Richard Wilson and Edmund Crouch; Risk and Culture, by Mary Douglas and Aaron Wildavsky. Bulletin of the Atomic Scientists, 39(6):33–38, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [DWH83].

Holdren:1983:NPN

[Hol83b] John P. Holdren. Nuclear power and nuclear weapons: the connection is dangerous. Bulletin of the Atomic Scientists, 39(1): 40–45, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Ano83c] and comments [FH83].

Holloway:1983:WTP

[Hol83c] David Holloway. The weapons tutorial — Part 4: Nuclear weapons in Europe. Bulletin of the Atomic Scientists, 39(4): 17–24, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Parts 1 [Mor83c], 2 [Tsi83i], 3 [Mar83b], 5 [Sar83], and 6 [Rom83].

Holdren:1984:PP

[Hol84a] John P. Holdren. Perspective: Pugwash 1984. Bulletin of the Atomic Scientists, 40(9):3–4, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hollingsworth:1984:SS

[Hol84b] Rogers Hollingsworth. The snare of specialization. Bulletin of the Atomic Scientists, 40(6):34–37, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holton:1984:MPU

[Hol84c] Gerald Holton. The migration of physicists to the United States. Bulletin of the Atomic Scientists, 40(4):18–24, April 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1985:NSI

[Hol85a] John P. Holdren. North–South issues and East–West confrontation. Bulletin of the Atomic Scientists, 41(7):97–101, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1985:IIL

[Hol85b] David Holloway. International issues: Lessons of the arms race. Bulletin of the Atomic Scientists, 41(7):80–83, August 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1986:BRBb

[Hol86a] David Holloway. Book review: Igor Vasil’evich Kurchatov [in the Physicotechnical] Institute, by A. P. Grinberg and V. Ia. Frenkel. Bulletin of the Atomic Scientists, 42(7):64, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1986:BRBa

[Hol86b] David Holloway. Book review: The Heavens and the Earth: A Political History of the Space Age, by Walter A. McDougall. Bulletin of the Atomic Scientists, 42(1):50–51, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1986:SPC

[Hol86c] David Holloway. The Soviet Party Congress. Bulletin of the Atomic Scientists, 42(5):15–19, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holst:1986:FTC

[Hol86d] John Jørgen Holst. Foreword: toward a common approach. Bulletin of the Atomic Scientists, 42(6):24–25, June/July 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1987:HB

[Hol87a] John P. Holdren. Harrison Brown 1917–1986. Bulletin of the Atomic Scientists, 43(2):3–8, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Owe87].

Holloway:1987:EMS

[Hol87b] David Holloway. [Excerpt: The making of the Soviet bomb]. Bulletin of the Atomic Scientists, 43(10):39, December 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1988:WGS

[Hol88] David Holloway. Why Gareyev’s statement is important. Bulletin of the Atomic Scientists, 44(10):32, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1990:MAS

[Hol90] David Holloway. In memoriam: Andrei Sakharov, 1921–1989. Bulletin of the Atomic Scientists, 46(2):14, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1991:BRM

[Hol91] David Holloway. Book reviews: Moral leader of a nation: Memoirs, by Andrei Sakharov; Moscow and Beyond, by Andrei Sakharov. Bulletin of the Atomic Scientists, 47(6):37–38, July/ August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1993:SSS

[Hol93] David Holloway. Soviet scientists speak out. Bulletin of the Atomic Scientists, 49(4):18–19, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1994:DS

[Hol94a] John P. Holdren. Dangerous surplus. Bulletin of the Atomic Scientists, 50(3):39–41, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1994:HBS

[Hol94b] David Holloway. How the bomb saved Soviet physics. Bulletin of the Atomic Scientists, 50(6):46–55, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gou95].

Holloway:1994:LCJ

[Hol94c] David Holloway. Letter: Caution justified. Bulletin of the Atomic Scientists, 50(1):62–63, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [KS93, Les93b].

Holman:1994:GCP

[Hol94d] Blan Holman. Giving Claiborne Parish the business. Bulletin of the Atomic Scientists, 50(3):53–57, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holman:1994:NC

[Hol94e] Blan Holman. Nuclear cheerleading. Bulletin of the Atomic Scientists, 50(3):56, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holum:1994:TAT

[Hol94f] John D. Holum. A treaty for all time. Bulletin of the Atomic Scientists, 50(6):19–22, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DE95].

Holt:1995:PML

[Hol95] Victoria K. Holt. Pay-more-later plan. Bulletin of the Atomic Scientists, 51(2):52–53, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1996:PBP

[Hol96a] John P. Holdren. Peace-building in the post-Cold War world. Bulletin of the Atomic Scientists, 52(2):29–32, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holloway:1996:IWI

[Hol96b] David Holloway. Introduction: The way it was. Bulletin of the Atomic Scientists, 52(6):53–54, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:1997:WR

[Hol97] John P. Holdren. Work with Russia. Bulletin of the Atomic Scientists, 53(2):42–45, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ram97].

Holl:2001:SC

[Hol01] Jack M. Holl. Staying in the club. Bulletin of the Atomic Scientists, 57(6):71–72, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/6/71.

Holloway:2003:PS

[Hol03] David Holloway. Political science. Bulletin of the Atomic Scientists, 59(1):69–70, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/69.

Holliday:2004:LEB

[Hol04] Bobbie Holliday. Letter to the Editor: Blast from the past. Bulletin of the Atomic Scientists, 60(6):5–??, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-124135855.
html.

Holmes:2005:ACRb

[Hol05a] James Holmes. Article commentary: Rough waters ahead. Bulletin of the Atomic Scientists, 61(6):24–25, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/6/24.

Holmes:2005:ACRa

[Hol05b] James R. Holmes. Article commentary: The right stuff. Bulletin of the Atomic Scientists, 61(5):25–26, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/5/25.

Holmes:2006:RW

[Hol06] James R. Holmes. The Roosevelt way. Bulletin of the Atomic Scientists, 62(3):19–20, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/3/19.

Holmes:2007:WIC

[Hol07] James R. Holmes. When interests collide. Bulletin of the Atomic Scientists, 63(3):21–23, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/3/21.

Holmes:2016:SCF

[Hol16] James R. Holmes. Sea changes: The future of nuclear deterrence. Bulletin of the Atomic Scientists, 72(4):228–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdren:2020:OCN

[Hol20] John P. Holdren. The overwhelming case for no first use. Bulletin of the Atomic Scientists, 76(1):3–7, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1701277.

Hoober:1963:LET

[Hoo63] Daniel Hoober. Letter to the Editor: Toward a new economic system. Bulletin of the Atomic Scientists, 19(2):34, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoopes:1981:CSS

[Hoo81] Townsend Hoopes. Commentary: Security with SALT. Bulletin of the Atomic Scientists, 37(6):60–61, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoodbhoy:1993:MBI

[Hoo93] Pervez Hoodbhoy. Myth building: the “Islamic” bomb. Bulletin of the Atomic Scientists, 49(5):42–49, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoodbhoy:2002:NG

[Hoo02] Pervez Hoodbhoy. Nuclear gamblers. Bulletin of the Atomic Scientists, 58(5):26–27, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/5/26.

Hoodbhoy:2020:NG

[Hoo20] Pervez Hoodbhoy. 2002: Nuclear gamblers. Bulletin of the Atomic Scientists, 76(6):410–411, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1847511.

Hopley:1949:PCD

[Hop49] Russell J. Hopley. Planning for civil defense. Bulletin of the Atomic Scientists, 5(4):111, 126, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hopkins:1977:WST

[Hop77] John C. Hopkins. Why not stop testing? a weapons designer’s view. Bulletin of the Atomic Scientists, 33(4):30–31, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hopper:1984:BND

[Hop84] Thomas Hopper. Breaking the nuclear deadlock: A call for defensive weapons. Bulletin of the Atomic Scientists, 40(5):54–56, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horn:1960:LED

[Hor60] E. C. Horn. Letter to the Editor: Defines his dislike. Bulletin of the Atomic Scientists, 16(3):113, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horowitz:1966:MSC

[Hor66] Irving Louis Horowitz. Michigan State and the CIA: A dilemma for social science. Bulletin of the Atomic Scientists, 22(7):26– 29, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horowitz:1968:ACS

[Hor68] Irving Louis Horowitz. The Americanization of conflict: Social science “Fiction” in action. Bulletin of the Atomic Scientists, 24(3):21–28, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horowitz:1971:SLM

[Hor71] Norman H. Horowitz. The search for life on Mars: Where we stand today. Bulletin of the Atomic Scientists, 27(9):13–17, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horowitz:1975:APM

[Hor75] Irving Louis Horowitz. America’s pursuit of the millennium. Bulletin of the Atomic Scientists, 31(3):22–28, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horowitz:1976:LMI

[Hor76] Wm. Horowitz. Letter: Man’s inhumanity. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Hornig:1977:BPS

[Hor77] Donald F. Hornig. Books: The President’s scientist: a private diary A Scientist at the White House. The Private Diary of President Eisenhower’s Special Assistant for Science and Technology, by George Kistiakowsky. Bulletin of the Atomic Scientists, 33(4): 62–64, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horan:1988:LSS

[Hor88a] Jack Horan. Latest scram at Savannah reactors. Bulletin of the Atomic Scientists, 44(10):7–8, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horan:1988:MWS

[Hor88b] Jack Horan. Making warheads: Savannah reactors: on line and in trouble. Bulletin of the Atomic Scientists, 44(1):24–28, January/ February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horelick:1988:PDO

[Hor88c] Arnold L. Horelick. Perspective: Détente by any other name. Bulletin of the Atomic Scientists, 44(8):5–6, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horan:1989:WNP

[Hor89] Jack Horan. World nuclear power operators unite. Bulletin of the Atomic Scientists, 45(3):39–40, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Horowitz:2014:CNM

[Hor14] Michael C. Horowitz. Coming next in military tech. Bulletin of the Atomic Scientists, 70(1):54–62, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/70/1/54.full.pdf+html.

Hoshima:1986:IH

[Hos86] Tsutomu Hoshima. The inheritance of Hiroshima. Bulletin of the Atomic Scientists, 42(6):37–38, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hostetter:1987:BRB

[Hos87a] Robert Hostetter. Book reviews: Nuclear Holocausts: Atomic War in Fiction, 1895–1984, by Paul Brians; Fictions of nuclear disaster, by David Dowling. Bulletin of the Atomic Scientists, 43 (10):45–46, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hostetter:1987:NHA

[Hos87b] Robert Hostetter. Nuclear Holocausts: Atomic War in Fiction, 1895–1984/Fictions of Nuclear Disaster (book). Bulletin of the Atomic Scientists, 43(10):45–46, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hotchkiss:1975:LEE

[Hot75] Willis L. Hotchkiss. Letters: Enough, enough, enough. Bulletin of the Atomic Scientists, 31(7):3, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hotchkiss:1976:LPH

[Hot76] Willis L. Hotchkiss. Letter: The power of an H-bomb. Bulletin of the Atomic Scientists, 32(1):2, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Houn:1972:CFP

[Hou72] Franklin W. Houn. Chinese foreign policy in perspective. Bulletin of the Atomic Scientists, 28(2):15–21, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Houck:1976:LSE

[Hou76] J. K. Houck. Letter: Solar energy economics. Bulletin of the Atomic Scientists, 32(4):3, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ben76b].

Hough:1982:SSP

[Hou82] Jerry F. Hough. Soviet succession and policy choices. Bulletin of the Atomic Scientists, 38(9):49–54, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hough:1985:SDM

[Hou85] Jerry F. Hough. Soviet decision-making on defense. Bulletin of the Atomic Scientists, 41(7):84–88, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Howe:1956:PVS

[How56] Quincy Howe. Part VI: Science, education, and communication: Science and the public. Bulletin of the Atomic Scientists, 12(9): 341–342, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Howard:1962:LAZ

[How62] Michael Howard. Limited armament zones in Europe. Bulletin of the Atomic Scientists, 18(2):9–14, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Howard:1963:LEA

[How63] Dagett H. Howard. Letter to the Editor: Aviation analogy. Bulletin of the Atomic Scientists, 19(2):36, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Howard:1965:BBE

[How65a] Michael Howard. Books: On Escalation: Metaphors and Scenarios, by Herman Kahn; An Introduction to Strategy, by André Beaufre. Bulletin of the Atomic Scientists, 21(8):25–27, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Howe:1965:BBT

[How65b] Quincy Howe. Books: Politics in the Twentieth Century, by Hans J. Morgenthau. Volume I: The Decline of American Politics; Volume II. The Impasse of American Foreign Policy; Volume III. The Restoration of American Policy. Bulletin of the Atomic Scientists, 21(4):36–37, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haas:1980:CRB

[HP80] Peter H. Haas and Dominic A. Paolucci. Commentary: On reviewing The War Game. Bulletin of the Atomic Scientists, 36(7): 59, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kar80].

Hibbs:2009:EL

[HP09] Mark Hibbs and Andreas Persbo. The ElBaradei legacy. Bulletin of the Atomic Scientists, 65(5):10–23, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/5/10.

Hickenlooper:1949:LFP

[HPL^{+}49] B. B. Hickenlooper, Sumner T. Pike, David Lilienthal, Detlev W. Bronk, A. N. Richards, and Alan Gregg. [Letters:] the fellowship program: Testimony before the Joint Committee. Bulletin of the Atomic Scientists, 5(6–7):166–176, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herman:1991:LWI

[HPR91] Edward S. Herman, David Peterson, and Ruth B. Russell. Letters: Was it just? was it war? Bulletin of the Atomic Scientists, 47(9):44–45, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fot91a, Lop91] and comment [Rus92a].

Hurwitz:1984:CLL

[HQL84] Henry Hurwitz, Jr., Catherine Thiel Quigg, and Robert W. Loftin. Commentary: Low-level radiation. Bulletin of the Atomic Scientists, 40(7):51–52, 54, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Qui84, Lof84].

Hoffmann:1951:LEI

[HR51] Banesh Hoffmann and Eugene Rabinowitch. Letter to the Editor: [impure science]. Bulletin of the Atomic Scientists, 7(9):287, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wyl51].

Hill:1981:PAS

[HRS81] David L. Hill, Eugene Rabinowitch, and John A. Simpson, Jr. 1945 perspective: The atomic scientists speak up. Bulletin of the Atomic Scientists, 37(1):23–25, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Life magazine, 29 October 1945.

Hesketh:1996:LUC

[HS96] R. V. Hesketh and Alexander R. Sich. Letters: Unnatural causes. Bulletin of the Atomic Scientists, 52(5):3, 61, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sic96c].

Hesketh:1997:LCL

[HS97] R. V. Hesketh and George S. Stanford. Letters: Chernobyl and the law of the conservation of energy. Bulletin of the Atomic Scientists, 53(6):65–66, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sta97].

Hassibi:2013:ESI

[HS13] Navid Hassibi and Tom Sauer. Easing sanctions on Iran might someday be necessary — but it won’t be easy. Bulletin of the Atomic Scientists, 69(5):46–55, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/5/46.

Hsieh:1965:SSN

[Hsi65] Alice Langley Hsieh. The Sino–Soviet nuclear dialogue1 1963. Bulletin of the Atomic Scientists, 21(1):16–21, January 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoffman:2018:MAT

[HSJH18] Robert R. Hoffman, Nadine Sarter, Matthew Johnson, and John K. Hawley. Myths of automation and their implications for military procurement. Bulletin of the Atomic Scientists, 74 (4):255–261, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1486615.

Hsu:1979:MSC

[Hsu79] Robert C. Hsu. Mass science in China. Bulletin of the Atomic Scientists, 35(2):27–30, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Haworth:2019:WDA

[HSV19] Alida R. Haworth, Scott D. Sagan, and Benjamin A. Valentino. What do Americans really think about conflict with nuclear North Korea? The answer is both reassuring and disturbing. Bulletin of the Atomic Scientists, 75(4):179–186, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1629576.

Hayes:2003:NNV

[HT03] Peter Hayes and Nina Tannenwald. Nixing nukes in Vietnam. Bulletin of the Atomic Scientists, 59(3):52–59, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/3/52.

Hua:2002:PUW

[Hua02] Vanessa Hua. Panama: U.S. weapons, U.S. mess? Bulletin of the Atomic Scientists, 58(4):18–20, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/4/18.

Hubbell:1957:LET

[Hub57] Ann Hubbell. Letter to the Editor: Too many people. Bulletin of the Atomic Scientists, 13(3):113, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hubbell:1960:LEM

[Hub60] Ann Hubbell. Letter to the Editor: Moving back the clock — and con. Bulletin of the Atomic Scientists, 16(3):80, March 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchins:1948:DSF

[HUC48] Robert M. Hutchins, Harold C. Urey, and Edward U. Condon. In defense of science and freedom — speeches at the Condon dinner. Bulletin of the Atomic Scientists, 4(6):173–175, June 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hudson:1960:LEE

[Hud60] J. F. Hudson. Letter to the Editor: Economic catastrophe. Bulletin of the Atomic Scientists, 16(4):145, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hudson:1970:CVG

[Hud70] J. F. Hudson. Communications: Very grave danger. Bulletin of the Atomic Scientists, 26(1):47, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hudson:1976:TMU

[Hud76] Richard Hudson. Time for mutations in the United Nations. Bulletin of the Atomic Scientists, 32(9):39–43, November 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Le 77].

Hudson:1977:ISL

[Hud77] Richard Hudson. The international struggle for a law of the sea. Bulletin of the Atomic Scientists, 33(10):14–20, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hudson:1978:CUN

[Hud78] Richard Hudson. Commentary: From the United Nations Special Session on Disarmament: ‘most useful ever’ or ‘elephant death dance’. Bulletin of the Atomic Scientists, 34(7):57–59, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano78j].

Huessy:1977:LSP

[Hue77] Hans Huessy. Letter: ‘soldiers for peace’. Bulletin of the Atomic Scientists, 33(9):2, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NB77, Kas77, Mac77a].

Huebner:2004:TP

[Hue04] Albert L. Huebner. Time to panic. Bulletin of the Atomic Scientists, 60(4):65–67, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/4/65.

Huebert:2012:AWC

[Hue12] Rob Huebert. Arctic 2030: What are the consequences of climate change?: The Canadian response. Bulletin of the Atomic Scientists, 68(4):17–21, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/4/17.

Hughes:1956:PAR

[Hug56] Donald J. Hughes. Positive aspects of the release of secret information. Bulletin of the Atomic Scientists, 12(5):169–172, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hughes:1957:BBW

[Hug57] Donald J. Hughes. Books: Atoms for the World, by Laura Fermi. Bulletin of the Atomic Scientists, 13(5):180, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hugenberger:1987:LPS

[Hug87] Scott Hugenberger. Letter: Persuading scientists. Bulletin of the Atomic Scientists, 43(1):59, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bev86].

Huges:1995:LSS

[Hug95] Liz Huges. Letter: Solar, sí; nuclear, no. Bulletin of the Atomic Scientists, 51(6):73, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi95a].

Hui:1987:IVS

[Hui87] Dou Hui. International views: A skeptical look at initiatives. Bulletin of the Atomic Scientists, 43(2):35–36, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hui:1991:NWO

[Hui91] Dou Hui. New world order: Order through peaceful coexistence. Bulletin of the Atomic Scientists, 47(5):24–25, June 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hull:1976:BOM

[Hul76] David L. Hull. Books: One man’s Joie de Vivre: Beast or Angel? Choices that Make Us Human, by René Dubos. Bulletin of the Atomic Scientists, 32(1):46–47, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hull:1978:BBL

[Hul78] David L. Hull. Books: The Limits of Altruism: An Ecologist’s View of Survival, by Garrett Hardin. Bulletin of the Atomic Scientists, 34(5):53–54, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hull:1998:BRP

[Hul98] Edmund J. Hull. Book review: Perils of peacekeeping: Blue Helmets: The Strategy of U.N. Military Operations, by John Hillen . Bulletin of the Atomic Scientists, 54(6):61–63, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1955:PCD

[Hum55] Hubert H. Humphrey. To provide for the common defense. Bulletin of the Atomic Scientists, 11(7):244–246, 268, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1958:USD

[Hum58] Hubert Humphrey. United States disarmament and foreign policy. Bulletin of the Atomic Scientists, 14(4):160, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1959:NSD

[Hum59a] Hubert H. Humphrey. New scientific data and test ban negotiations. Bulletin of the Atomic Scientists, 15(3):109–113, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1959:SR

[Hum59b] Hubert H. Humphrey. Senate Resolution 96. Bulletin of the Atomic Scientists, 15(5):186, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1960:SWR

[Hum60] Senator Hubert H. Humphrey. Should we resume nuclear weapons tests? Bulletin of the Atomic Scientists, 16(3):85–86, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1966:IC

[Hum66] Hubert H. Humphrey. [Interdependence and cooperation]. Bulletin of the Atomic Scientists, 22(1):3, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1972:IAL

[Hum72] Hubert H. Humphrey. Issues in arms limitations — four views: II. An end to nuclear gamesmanship. Bulletin of the Atomic Scientists, 28(3):12–15, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Humphrey:1977:CPD

[Hum77] J. M. Humphrey. The challenge of parasitic diseases. Bulletin of the Atomic Scientists, 33(3):46–53, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hummert:1983:CNP

[Hum83] Vic Hummert. Commentary: Nuclear power near Hong Kong. Bulletin of the Atomic Scientists, 39(3):62–63, March 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi82a].

Hunt:1984:CDN

[Hun84] Charles B. Hunt. Commentary: Disposal of nuclear wastes. Bulletin of the Atomic Scientists, 40(4):44–45, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Law84a].

Hunt:1985:BRB

[Hun85a] Charles B. Hunt. Book review: The Atom and the Fault, by Richard L. Meehan. Bulletin of the Atomic Scientists, 41(2):48– 49, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hunt:1985:USC

[Hun85b] Linda Hunt. United States cover-up of Nazi scientists. Bulletin of the Atomic Scientists, 41(4):16–24, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [OH85].

Huntington:1985:LPR

[Hun85c] W. R. Huntington. Letter: Peace requires world law. Bulletin of the Atomic Scientists, 41(3):62, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nag84].

Hune:1987:BRB

[Hun87] Shirley Hune. Book reviews: Third World Coups d’État and International Security, by Steven R. David; The Challenge to U.S. Policy in the Third World: Regional Responsibilities and Regional Devolution, by Thomas Perry Thornton. Bulletin of the Atomic Scientists, 43(9):47–49, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hunt:1990:LBN

[Hun90] Helen M. Hunt. Letter: Ban naval cruise and end nuclear trespass. Bulletin of the Atomic Scientists, 46(2):45, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coc89].

Huntley:2000:TMM

[Hun00] Wade Huntley. Too much muscle. Bulletin of the Atomic Scientists, 56(6):71–74, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/6/71.

Huntley:2003:ST

[Hun03] Wade L. Huntley. Sit down and talk. Bulletin of the Atomic Scientists, 59(4):28–29, July/August 2003. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/4/28.

Hunner:2004:IAG

[Hun04] Jon Hunner. Inventing Los Alamos: the growth of an atomic community. University of Oklahoma Press, Norman, OK, USA, 2004. ISBN 0-8061-3634-0. x + 288 pp. LCCN F804.L6 H86 2004. URL

http://www.loc.gov/catdir/toc/fy0604/2004046086.html.

Hunger:2014:WBA

[Hun14] Iris Hunger. Winning the battle against emerging pathogens: a German response. Bulletin of the Atomic Scientists, 70(4):22–25, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Huq:2011:LCC

[Huq11] Saleemul Huq. Lessons of climate change, stories of solutions: Bangladesh: Adaptation. Bulletin of the Atomic Scientists, 67(1): 56–59, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/67/1/56.

Huq:2016:CCT

[Huq16] Saleemul Huq. Climate change, technological innovation: a Bangladeshi response. Bulletin of the Atomic Scientists, 72(1): 2–3, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hurd:1960:CD

[Hur60] Douglas Hurd. A case for the diplomats. Bulletin of the Atomic Scientists, 16(2):52–53, 57, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hurwitz:1974:LMF

[Hur74] Henry Hurwitz. Letter: Moratorium folly. Bulletin of the Atomic Scientists, 30(8):2, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel74b, Mic74b].

Hurwitz:1975:LOW

[Hur75] Henry Hurwitz. Letters: The only way. Bulletin of the Atomic Scientists, 31(8):3, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com74].

Hurwitz:1977:LI

[Hur77a] Henry Hurwitz, Jr. Letter: An inconsistency. Bulletin of the Atomic Scientists, 33(1):6–7, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds76].

Hurwitz:1977:LPW

[Hur77b] Henry Hurwitz, Jr. Letter: Planning for the worst. Bulletin of the Atomic Scientists, 33(5):4, 6, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH77b].

Hurley:1978:BBW

[Hur78a] John G. Hurley. Books: Scientists and World Order: The Uses of Technical Knowledge in International Organization, by Ernst B. Haas, Mary Pat Williams, and Don Babai. Bulletin of the Atomic Scientists, 34(7):49, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hurwitz:1978:LRW

[Hur78b] Henry Hurwitz. Letter: Real world solutions. Bulletin of the Atomic Scientists, 34(5):5, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol78, Wei78f, HWH78].

Hurwitz:1988:LSM

[Hur88] Bruce Hurwitz. Letter: Seven minutes to midnight? Bulletin of the Atomic Scientists, 44(4):53, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack88h].

Husted:1953:SAA

[Hus53] Ellery Husted. Shelter in the Atomic Age. Bulletin of the Atomic Scientists, 9(7):273–276, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Husbands:1990:BMA

[Hus90] Jo L. Husbands. A buyer’s market for arms. Bulletin of the Atomic Scientists, 46(4):14–16, 17–19, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Husain:2003:NMB

[Hus03] Khurram Husain. Neocons: The men behind the curtain. Bulletin of the Atomic Scientists, 59(6):62–71, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/6/62.

Husain:2004:PDP

[Hus04] Khurram Husain. Pakistan: Disparities in power. Bulletin of the Atomic Scientists, 60(2):24–26, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/24.

Huss:2024:AIG

[Hus24] Matthias Huss. The Alps’ iconic glaciers are melting, but there’s still time to save the biggest. Bulletin of the Atomic Scientists, 80(4):225–229, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2364501.

Hutchins:1946:PWR

[Hut46a] Robert M. Hutchins. Peace or war with Russia? Bulletin of the Atomic Scientists of Chicago, 1(6):1–2, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchinson:1946:MPD

[Hut46b] W. S. Hutchinson. The Manhattan Project declassification program. Bulletin of the Atomic Scientists, 2(9–10):14–15, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchinson:1950:CDN

[Hut50] Clyde A. Hutchinson, Jr. Civil defense news. Bulletin of the Atomic Scientists, 6(12):382, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchinson:1951:CDN

[Hut51] Clyde A. Hutchinson, Jr. Civil defense news: Progress in the states. Bulletin of the Atomic Scientists, 7(3):94, 96, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchins:1953:SPC

[Hut53a] Robert M. Hutchins. Social and political conformity. Bulletin of the Atomic Scientists, 9(4):105–108, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchins:1953:WFI

[Hut53b] Robert M. Hutchins. “The way to fight ideas”. Bulletin of the Atomic Scientists, 9(5):192, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hutchins:1961:BBV

[Hut61] Robert M. Hutchins. Books: The Voice of the Dolphins, by Leo Szilard. Bulletin of the Atomic Scientists, 17(7):290, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Pag61].

Huxley:1947:RRP

[Hux47] Aldous Huxley. Replies to Ridenour: A positive program of research for peace. Bulletin of the Atomic Scientists, 3(8):225, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Huxley:1949:FSA

[Hux49] Julian Huxley. Freedom for science: An appeal for action. Bulletin of the Atomic Scientists, 5(8–9):209–210, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Huxley:1951:TS

[Hux51] Julian Huxley. Taking stock. Bulletin of the Atomic Scientists, 7(7–8):209, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Huxley:1959:C

[Hux59a] Julian Huxley. A comment. Bulletin of the Atomic Scientists, 15(9):375, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Remark supporting the theory of evolution.

Huxley:1959:FM

[Hux59b] Julian Huxley. The future of man. Bulletin of the Atomic Scientists, 15(10):402–404, 409, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hoffman:2018:INT

[HV18] Wyatt Hoffman and Tristan A. Volpe. Internet of nuclear things: Managing the proliferation risks of 3-D printing technology. Bulletin of the Atomic Scientists, 74(2):102–113, 2018.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1436811.

Hall:1975:LMT

[HvH75] William Cornelius Hall and Frank von Hippel. Letters: Make them better. Bulletin of the Atomic Scientists, 31(8):2, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [PvH74].

Hayes:2006:GL

[HVK^{+}06] Peter Hayes, David Von Hippel, Jungmin Kang, Tatsujiro Suzuki, Richard Tanter, and Scott Bruce. Grid-locked. Bulletin of the Atomic Scientists, 62(1):52–58, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/52.

Hammond:1973:CEL

[HW73] R. Philip Hammond and Alvin M. Weinberg. Communications: Energy limits. Bulletin of the Atomic Scientists, 29(5):2–3, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holdstock:1996:LEN

[HW96] Douglas Holdstock and Liz Waterston. Letter: A European NWFZ? Bulletin of the Atomic Scientists, 52(4):3, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [FC96].

Hodge:2008:ERN

[HW08] Nathan Hodge and Sharon Weinberger. The ever-ready nuclear missileer. Bulletin of the Atomic Scientists, 64(3):14–21, July/ August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/64/
3/14.

Hallion:1995:LAS

[HWG95] Richard P. Hallion, Herman S. Wolk, and Stanley Goldberg. Letter: Air and Space Museum guilty, as charged. Bulletin of the Atomic Scientists, 51(4):75–76, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol95c].

Holdren:1978:LRS

[HWH78] John P. Holdren, Leonard Weiss, and Henry Hurwitz, Jr. Letter: Un-real solutions. Bulletin of the Atomic Scientists, 34(6):2–3, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hur78b].

Horelick:1967:CAR

[HY67] Arnold Horelick and Elizabeth Young. Comments: The author and the reviewer. Bulletin of the Atomic Scientists, 23(5):62–63, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hyde:1975:BBB

[Hyd75] Steven Hyde. Books: Sixth Book of Mathematical Games From Scientific American, by Martin Gardner. Bulletin of the Atomic Scientists, 31(8):48, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hyder:1991:LBR

[Hyd91] James Hyder. Letter: B Reactor no place for tourists. Bulletin of the Atomic Scientists, 47(1):46, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sto90].

Hymes:1968:BBD

[Hym68] Dell Hymes. Books: The Dissenting Academy, by Theodore Raszak. Bulletin of the Atomic Scientists, 24(9):29–34, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ros69, Hym69b].

Hyman:1969:MMC

[Hym69a] Sidney Hyman. Man on the Moon — the Columbian dilemma. Bulletin of the Atomic Scientists, 25(7):17–22, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hymes:1969:CHR

[Hym69b] Dell Hymes. Communications: Hymes replies. Bulletin of the Atomic Scientists, 25(8):47, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hym68, Ros69].

Hymoff:1971:SIC

[Hym71] Edward Hymoff. Stalemate in Indo-China: Technology vs. guerrillas. Bulletin of the Atomic Scientists, 27(9):27–30, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hyman:1991:BRC

[Hym91] Sidney Hyman. Book review: Clifford’s brief: Counsel to the President, by Clark Clifford. Bulletin of the Atomic Scientists, 47(8):40–41, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hymans:2007:NKN

[Hym07] Jacques E. C. Hymans. North Korea’s nuclear neurosis. Bulletin of the Atomic Scientists, 63(3):??, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/3/44.

Hynek:1969:BBS

[Hyn69] J. Allen Hynek. Books: Scientific Study of Unidentified Flying Objects, by Edward U. Condon. Bulletin of the Atomic Scientists, 25(4):39–42, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ialenti:2018:WMH

[Ial18] Vincent Ialenti. Waste makes haste: How a campaign to speed up nuclear waste shipments shut down the WIPP long-term repository. Bulletin of the Atomic Scientists, 74(4):262–275, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1486616.

Ibrahim:2013:ACW

[Ibr13] Jamal Khaer Ibrahim. Attracting a crowd: What societal verification means for arms control : The Malaysian response. Bulletin of the Atomic Scientists, 69(3):19–22, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/3/19.

Ibser:1987:LYP

[Ibs87] H. W. Ibser. Letter: Yield, not power. Bulletin of the Atomic Scientists, 43(8):54, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NA87a].

Ichheiser:1964:LEN

[Ich64] Gustav Ichheiser. Letter to the Editor: Is nationalism really outmoded? Bulletin of the Atomic Scientists, 20(6):35, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ide:1993:LTW

[Ide93] Hiroshi Ide. Letter: They weren’t “tests”. Bulletin of the Atomic Scientists, 49(6):54, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NA93h].

Idso:1981:CCP

[Ids81] Sherwood B. Idso. Commentary: On CO_{2}: Proceed with caution’. Bulletin of the Atomic Scientists, 37(7):62, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bar81d].

Inglis:1979:CCE

[IES79] David Rittenhouse Inglis, Charles E. Engelke, and Leo Seren. Commentary: Community energy systems. Bulletin of the Atomic Scientists, 35(4):50–52, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eng78].

Inglis:1951:DBM

[IF51] David R. Inglis and Donald A. Flanders. A deal before midnight? Bulletin of the Atomic Scientists, 7(10):305–306, 317, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ifft:1967:RSS

[Iff67] Edward M. Ifft. Reports: Science students at Moscow U. Bulletin of the Atomic Scientists, 23(4):37–40, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BCA

[IG91] John Isaacs and Joe Goldman. Bulletin: Closing the arms bazaar. Bulletin of the Atomic Scientists, 47(4):4–5, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:2001:TPF

[IK01] John Isaacs and Dan Koslofsky. Trim Pentagon fat. Bulletin of the Atomic Scientists, 57(1):26–28, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/26.

Ifft:2021:LPM

[IK21] Edward M. Ifft and David A. Koplow. Legal and political myths of the Treaty on the Prohibition of Nuclear Weapons. Bulletin of the Atomic Scientists, 77(3):134–139, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1912198.

Ikegami:2016:SPT

[Ike16] Masako Ikegami. Seeking a path toward missile nonproliferation. Bulletin of the Atomic Scientists, 72(6):365–367, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ikle:1960:TCD

[Ikl60] Fred Charles Iklé. N th countries and disarmament. Bulletin of the Atomic Scientists, 16(10):391–394, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ikle:1975:NWN

[Ikl75] Fred C. Iklé. The nether world of nuclear megatonnage. Bulletin of the Atomic Scientists, 31(1):20–24, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ikle:1976:IRA

[Ikl76] Fred C. Iklé. Illusions and realities about nuclear energy. Bulletin of the Atomic Scientists, 32(8):14–17, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ikle:1980:BRN

[Ikl80] Fred Charles Iklé. Bombs and reactors: the nuclear divide. Bulletin of the Atomic Scientists, 36(1):38–42, January 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Iltis:1970:COH

[ILA70] Hugh H. Iltis, Orie L. Loucks, and Peter Andrews. Criteria for an optimum human environment. Bulletin of the Atomic Scientists, 26(1):2–6, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ilnitsky:1997:RRM

[Iln97] Andrei Ilnitsky. Report: Russia: Mysteries under Moscow. Bulletin of the Atomic Scientists, 53(3):11–14, May/June 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ilnitsky:1997:BKC

[IM97] Andred Ilnitsky and Nancy Myers. Bulletin: The Kremlin capsule. Bulletin of the Atomic Scientists, 53(2):7–8, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moon:2016:NKN

[iM16] Chung in Moon. North Korea, nuclear weapons, and the search for a new path forward. Bulletin of the Atomic Scientists, 72(5): 343–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Imai:1969:NPT

[Ima69] Ryukichi Imai. The Non-Proliferation Treaty and Japan. Bulletin of the Atomic Scientists, 25(5):2–7, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Imai:1970:PIA

[Ima70] Ryukichi Imai. Part 2: International atom: Japan and the nuclear age. Bulletin of the Atomic Scientists, 26(6):35–39, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Imai:1971:JWS

[Ima71] Ryukichi Imai. Japan and the world of SALT. Bulletin of the Atomic Scientists, 27(10):13–16, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Imai:1993:AAR

[Ima93a] Ryukichi Imai. Asian ambitions, rising tensions. Bulletin of the Atomic Scientists, 49(5):33–36, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Imai:1993:FPN

[Ima93b] Ryukichi Imai. A fine point of nuclear theology. Bulletin of the Atomic Scientists, 49(5):34, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Imshenetsky:1960:MMB

[Ims60] A. A. Imshenetsky. Modern microbiology and the biological warfare menace. Bulletin of the Atomic Scientists, 16(6):241–242, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ICDSI:1981:GVS

[Ind81] Independent Commission on Disarmament and Security Issues. Global voices for SALT. Bulletin of the Atomic Scientists, 37(7): 6, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Accident:2014:PC

[Ind14] Independent Investigation Commission on the Fukushima Nuclear Accident and Masaharu Fujiyoshi. Prologue to catastrophe. Bulletin of the Atomic Scientists, 70(2):36–41, March 2014. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Infeld:1962:FS

[Inf62] Leopold Infeld. Freedom for science. Bulletin of the Atomic Scientists, 18(5):31, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Infeld:1965:SI

[Inf65] Léopold Infeld. As I see it. Bulletin of the Atomic Scientists, 21(2):7–14, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Infeld:2006:QA

[Inf06] Leopold Infeld. Quest: an autobiography. AMS Chelsea Publishing, Providence, RI, USA, second edition, 2006. ISBN 0-8218-4073-8, 0-8284-0309-0. 361 pp. LCCN QC16.I6 A3 1980.

Inglis:1946:AVP

[Ing46] David R. Inglis. The ADA and the veto power. Bulletin of the Atomic Scientists, 1(12):15, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1952:NWP

[Ing52a] David R. Inglis. The new working paper and U.S. atomic policy. Bulletin of the Atomic Scientists, 8(5):132–134, 138, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1952:TAW

[Ing52b] David R. Inglis. Tactical atomic weapons and the problem of ultimate control. Bulletin of the Atomic Scientists, 8(3):79–83, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1952:DCP

[Ing52c] David R. Inglis. UN Disarmament Commission proceedings. Bulletin of the Atomic Scientists, 8(8):271–274, November 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1953:APQ

[Ing53] David R. Inglis. Atomic profits and the question of survival. Bulletin of the Atomic Scientists, 9(4):114–119, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1954:BHB

[Ing54a] David R. Inglis. Ban H-bomb tests and favor the defense. Bulletin of the Atomic Scientists, 10(9):353–356, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1954:HBD

[Ing54b] David R. Inglis. The H-bomb and disarmament prospects. Bulletin of the Atomic Scientists, 10(2):41–45, 64, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1955:GEW

[Ing55a] David R. Inglis. Guest editorial: We haven’t really tried. Bulletin of the Atomic Scientists, 11(1):3–4, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1955:SAT

[Ing55b] David R. Inglis. The Stassen appointment: Turning point in disarmament thinking? Bulletin of the Atomic Scientists, 11(6): 204–205, 216, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1956:PIS

[Ing56a] David Inglis. Part II: Science and military strategy: National security with the arms race limited. Bulletin of the Atomic Scientists, 12(6):196–201, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1956:ADD

[Ing56b] David R. Inglis. Armament decision in a democracy. Bulletin of the Atomic Scientists, 12(8):284–286, 319, October 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1956:BBB

[Ing56c] David R. Inglis. Books: The Bomb — Challenge and Answer, by Alexander Haddow, Bertrand Russell, Lord [William Henry] Beveridge, and Henry Usborne. Bulletin of the Atomic Scientists, 12(5):184, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1957:ACE

[Ing57a] David R. Inglis. Arms control effort buried in State [Department]. Bulletin of the Atomic Scientists, 13(5):174–175, May 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1957:PSN

[Ing57b] David R. Inglis. Prospects for stopping nuclear tests. Bulletin of the Atomic Scientists, 13(1):19–23, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1959:AST

[Ing59a] David R. Inglis. Allaying suspicions of test ban controls. Bulletin of the Atomic Scientists, 15(10):425–426, December 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1959:FCP

[Ing59b] David R. Inglis. The fourth-country problem: Let’s stop at three. Bulletin of the Atomic Scientists, 15(1):22–26, January 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1960:CHT

[Ing60a] David R. Inglis. The Congressional hearings on technical aspects of test control. Bulletin of the Atomic Scientists, 16(6):205–207, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1960:EFT

[Ing60b] David R. Inglis. Excessive fear of test ban evasion. Bulletin of the Atomic Scientists, 16(5):168–171, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1962:SCW

[Ing62] David R. Inglis. Shelters and the chance of war. Bulletin of the Atomic Scientists, 18(4):18–21, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1963:DAC

[Ing63a] David R. Inglis. Disarmament after Cuba. Bulletin of the Atomic Scientists, 19(1):18–21, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1963:RTB

[Ing63b] David R. Inglis. The rest of the test ban. Bulletin of the Atomic Scientists, 19(10):39–40, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1964:LET

[Ing64] David R. Inglis. Letter to the Editor: The test ban treaty. Bulletin of the Atomic Scientists, 20(1):26, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1965:EOC

[Ing65] David R. Inglis. The explosion of October 16: The Chinese bombshell. Bulletin of the Atomic Scientists, 21(2):19–21, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1967:MDN

[Ing67] David R. Inglis. Missile defense, nuclear spread, and Vietnam. Bulletin of the Atomic Scientists, 23(5):49–52, May 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1968:CJM

[Ing68a] David R. Inglis. Conservative judgments and missile madness. Bulletin of the Atomic Scientists, 24(5):6–11, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1968:NTA

[Ing68b] David R. Inglis. Nuclear threats, ABM systems, and proliferation. Bulletin of the Atomic Scientists, 24(6):2–4, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1970:BBC

[Ing70] David Rittenhouse Inglis. Books: The Careless Atom, by Sheldon Novick; The Perils of the Peaceful Atom, by Richard Curtis and Elizabeth Hogan. Bulletin of the Atomic Scientists, 26(2):50– 54, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [LI70].

Inglis:1971:NEM

[Ing71] David Rittenhouse Inglis. Nuclear energy and the Malthusian dilemma. Bulletin of the Atomic Scientists, 27(2):14–18, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WI71].

Inglis:1974:GES

[Ing74] David Rittenhouse Inglis. Guest editorial: The sweet voice of reason. Bulletin of the Atomic Scientists, 30(7):50–52, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1975:WPN

[Ing75] David Rittenhouse Inglis. Wind power now! Bulletin of the Atomic Scientists, 31(8):20–26, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zel76a] and reply [Ing76].

Inglis:1976:PMS

[Ing76] David Rittenhouse Inglis. The potential must be stressed to get the bureaucracy moving. Bulletin of the Atomic Scientists, 32 (3):60, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ing75, Zel76a].

Inglis:1977:LMT

[Ing77] David Rittenhouse Inglis. Letter: Moving too slowly. Bulletin of the Atomic Scientists, 33(8):2, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1979:BRBb

[Ing79a] David Rittenhouse Inglis. Book review: Offshore and Underground Power Plants, edited by Robert Noyes. Bulletin of the Atomic Scientists, 35(5):36–38, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1979:BRBa

[Ing79b] David Rittenhouse Inglis. Book reviews: Blowing on the Wind: The Nuclear Test Ban Debate, 1954–1960, by Robert A. Devine. Bulletin of the Atomic Scientists, 35(2):45–46, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ing:1982:BRB

[Ing82] Stanley C. M. Ing. Book review: New Directions in Disarmament, edited by William Epstein and Bernard T. Feld. Bulletin of the Atomic Scientists, 38(4):42–43, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1984:BND

[Ing84] David Rittenhouse Inglis. Breaking the nuclear deadlock: Freeze the cruise. Bulletin of the Atomic Scientists, 40(1):48–50, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1985:MDM

[Ing85] David Rittenhouse Inglis. Minimum deterrence, maximum stability. Bulletin of the Atomic Scientists, 41(3):42, 46–49, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ICRC:1950:RCA

[Int50] International Committee of the Red Cross. Red Cross appeal to governments for prohibition of atomic weapons. Bulletin of the Atomic Scientists, 6(7):223–224, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

IPCC:1967:PXC

[Int67] International Pugwash Continuing Committee. Pugwash XVII [Conference on Science and World Affairs, Ronneby, Sweden, 3–8 September 1967]. Bulletin of the Atomic Scientists, 23(9):46– 48, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ICPM:1978:CRN

[Int78] International Council of the Pugwash Movement. A convention for the renunciation of nuclear weapons. Bulletin of the Atomic Scientists, 34(5):43–44, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

IPPNW:1981:CWW

[Int81] International Physicians for the Prevention of Nuclear War. Commentary: What we can do. Bulletin of the Atomic Scientists, 37 (6):20–21, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ITFPNT:1986:RNT

[Int86a] International Task Force on Prevention of Nuclear Terrorism. Report on nuclear terrorism. Bulletin of the Atomic Scientists, 42 (10):38–42, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ITFPNT:1986:TFR

[Int86b] International Task Force on Prevention of Nuclear Terrorism. Task force recommendations. Bulletin of the Atomic Scientists, 42(10):42–44, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kri87].

IPCC:1992:AC

[Int92] Intergovernmental Panel on Climate Change. Ambiguous conclusion. Bulletin of the Atomic Scientists, 48(5):38, June 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ince:1984:CPC

[IPB84] Martin Ince, George Palmer, and Dan I. Bolef. Commentary: On the plutonium connection. Bulletin of the Atomic Scientists, 40(6):56, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [PB84].

Irion:1956:SNF

[Iri56] Frederick C. Irion. Sir Nuclear Fission: A realistic myth. Bulletin of the Atomic Scientists, 12(5):179–180, 183, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Pri55].

Iriye:1979:BRB

[Iri79] Akira Iriye. Book review: China and the Major Powers in East Asia, by A. Doak Barnett. Bulletin of the Atomic Scientists, 35 (5):39–40, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Iriye:1986:BRB

[Iri86] Akira Iriye. Book review: Contending with Kennan, by Barton Gellman. Bulletin of the Atomic Scientists, 42(8):52–53, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Inglis:1967:SRP

[IS67] David R. Inglis and Carl L. Sandler. A special report on Plowshare: Prospects and problems: The nonmilitary uses of nuclear explosives. Bulletin of the Atomic Scientists, 23(10):46–53, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Iakimets:1992:PNT

[IS92] Vladimir Iakimets and Olzhas Suleimenov. Perspective: New tests mean new nukes. Bulletin of the Atomic Scientists, 48(8): 12–13, 45, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Islam:2015:UWC

[IS15] Shafiqul Islam and Lawrence Susskind. Understanding the water crisis in Africa and the Middle East: How can science inform policy and practice? Bulletin of the Atomic Scientists, 71(2): 39–49, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1978:WSI

[Isa78] John Isaacs. From Washington: The SALT II campaign begins. Bulletin of the Atomic Scientists, 34(7):60–61, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1980:FMP

[Isa80] John David Isaacs. The FY 1981 military programs. Bulletin of the Atomic Scientists, 36(6):34–40, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1981:CWT

[Isa81a] John D. Isaacs. Commentary from Washington: Three R’s of arms control: Reagan, Rostow and Rowny. Bulletin of the Atomic Scientists, 37(7):5–7, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1981:CWR

[Isa81b] John David Isaacs. Commentary from Washington: Reagan’s ‘defense’ budget. Bulletin of the Atomic Scientists, 37(5):12–13, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1982:MB

[Isa82a] John Isaacs. 1983 military budget. Bulletin of the Atomic Scientists, 38(4):16–18, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1982:EF

[Isa82b] John Isaacs. Election 1982: The freeze. Bulletin of the Atomic Scientists, 38(8):9–11, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1982:SW

[Isa82c] John Isaacs. The summer in Washington. Bulletin of the Atomic Scientists, 38(7):8–10, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1982:DBL

[Isa82d] John D. Isaacs. Defense budget: look but don’t touch. Bulletin of the Atomic Scientists, 38(2):12–15, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1983:WNA

[Isa83a] John Isaacs. From Washington: Nervous about nerve gas. Bulletin of the Atomic Scientists, 39(10):7–8, December 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pay84].

Isaacs:1983:WHN

[Isa83b] John Isaacs. What happened to the Ninety-eighth Congress. Bulletin of the Atomic Scientists, 39(7):11–12, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1983:BLM

[Isa83c] John D. Isaacs. Bulletin: The lobbyist and the MX. Bulletin of the Atomic Scientists, 39(2):56–57, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1984:DBE

[Isa84a] John Isaacs. The defense budget in an election year. Bulletin of the Atomic Scientists, 40(2):3–4, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gro84a].

Isaacs:1984:WMB

[Isa84b] John Isaacs. From Washington: The military budget goes to Congress. Bulletin of the Atomic Scientists, 40(4):7–8, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1985:OCA

[Isa85a] John Isaacs. 1985 outlook: Congress and the arms control paradox. Bulletin of the Atomic Scientists, 41(1):9–10, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1985:WRM

[Isa85b] John Isaacs. Washington report: MX: Reagan’s pyrrhic victory? Bulletin of the Atomic Scientists, 41(6):4–6, June/July 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1985:WRNa

[Isa85c] John Isaacs. Washington report: New lineup on defense committees. Bulletin of the Atomic Scientists, 41(3):7–8, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1985:WRNb

[Isa85d] John Isaacs. Washington report: November — critical month for arms control. Bulletin of the Atomic Scientists, 41(8):3–4, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1985:WRU

[Isa85e] John Isaacs. Washington report: Use and abuse of Geneva talks. Bulletin of the Atomic Scientists, 41(5):4–5, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1985:WRW

[Isa85f] John Isaacs. Washington report: Will Aspin sink or swim? Bulletin of the Atomic Scientists, 41(10):8–9, November 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1986:WRCa

[Isa86a] John Isaacs. Washington report: Committees key to arms decisions. Bulletin of the Atomic Scientists, 42(1):6–8, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1986:WRCb

[Isa86b] John Isaacs. Washington report: Congress and the military revisited. Bulletin of the Atomic Scientists, 42(2):6–7, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1986:WRH

[Isa86c] John Isaacs. Washington report: House challenges Reagan on arms control. Bulletin of the Atomic Scientists, 42(8):6–7, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1986:WRR

[Isa86d] John Isaacs. Washington report: Reagan wins as loses. Bulletin of the Atomic Scientists, 42(6):6–7, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1986:WRS

[Isa86e] John Isaacs. Washington report: Showdown on military budget. Bulletin of the Atomic Scientists, 42(4):3–4, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1986:WRU

[Isa86f] John Isaacs. Washington report: Using summitry to thwart Congress. Bulletin of the Atomic Scientists, 42(10):4–5, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1987:ACN

[Isa87a] John Isaacs. Arms control in the new Congress. Bulletin of the Atomic Scientists, 43(1):4–5, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1987:SRA

[Isa87b] John Isaacs. Senate Republican arms control lineup. Bulletin of the Atomic Scientists, 43(7):6, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1987:WRC

[Isa87c] John Isaacs. Washington report: Congress tries again on arms control. Bulletin of the Atomic Scientists, 43(5):3–4, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1987:WRSb

[Isa87d] John Isaacs. Washington report: Senate infighting on treaty. Bulletin of the Atomic Scientists, 43(9):3–4, November 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1987:WRSa

[Isa87e] John Isaacs. Washington report: Senate minority could imperil treaty. Bulletin of the Atomic Scientists, 43(7):5–6, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1987:WRF

[Isa87f] John Isaacs. Washington report: The fall and rise of Les Aspin. Bulletin of the Atomic Scientists, 43(3):4–5, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1988:WRP

[Isa88a] John Isaacs. Washington report: Pressing deadlines; transition fever; who would thunk it? Bulletin of the Atomic Scientists, 44(10):3–4, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1988:WRRa

[Isa88b] John Isaacs. Washington report: Reagan’s arms control record. Bulletin of the Atomic Scientists, 44(1):4–5, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cob88].

Isaacs:1988:WRS

[Isa88c] John Isaacs. Washington report: Star Wars vote zapped; high priorities; odd couple fights depression; the world according to Gaffney; House yeas, Senate nays on arms control. Bulletin of the Atomic Scientists, 44(6):3–4, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1988:WRRb

[Isa88d] John Isaacs. Washington report: The ratification circus. Bulletin of the Atomic Scientists, 44(3):3–4, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1988:WRWb

[Isa88e] John Isaacs. Washington report: War of the one-liners; does hawk loss mean dove win? stopping the nonsense. Bulletin of the Atomic Scientists, 44(8):3–4, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1988:WRWa

[Isa88f] John Isaacs. Washington report: What a difference a decade makes; Jesse Helms tries constructive engagement; Midgetman’s last gasps; curbing chemical dependency. Bulletin of the Atomic Scientists, 44(5):3–4, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BDV

[Isa89a] John Isaacs. Bulletin: Discord, verse two. Bulletin of the Atomic Scientists, 45(10):6, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BDH

[Isa89b] John Isaacs. Bulletin: Dueling headlines. Bulletin of the Atomic Scientists, 45(8):6, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BNS

[Isa89c] John Isaacs. Bulletin: Nunn’s SDI two-step. Bulletin of the Atomic Scientists, 45(10):5–6, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BTN

[Isa89d] John Isaacs. Bulletin: Time for a new speechwriter. Bulletin of the Atomic Scientists, 45(9):5–6, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BVV

[Isa89e] John Isaacs. Bulletin: Voice in the vacuum. Bulletin of the Atomic Scientists, 45(8):6–7, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BWD

[Isa89f] John Isaacs. Bulletin: Where does Bush stand on chemical weapons? Bulletin of the Atomic Scientists, 45(10):3, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:BWM

[Isa89g] John Isaacs. Bulletin: Who’s micromanaging the Pentagon? Bulletin of the Atomic Scientists, 45(8):6, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:DBM

[Isa89h] John Isaacs. Debate at $3 billion a minute. Bulletin of the Atomic Scientists, 45(9):3–4, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:WRB

[Isa89i] John Isaacs. Washington report: B-2 or not B-2?; one MXed-up debate; less than meets eye-in-the-sky. Bulletin of the Atomic Scientists, 45(7):3–4, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:WRG

[Isa89j] John Isaacs. Washington report: George Bush moves in. Bulletin of the Atomic Scientists, 45(1):3–4, January/February 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano90-53].

Isaacs:1989:WRL

[Isa89k] John Isaacs. Washington report: Let’s make a deal; hawks lose prey; doves’ golden silence; learning to love draft-dodgers. Bulletin of the Atomic Scientists, 45(5):3–4, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:WRN

[Isa89l] John Isaacs. Washington report: No battleships in the Bay; what price machismo?; the general’s little missile; they gave a budget and no one came; clash of the dogmas. Bulletin of the Atomic Scientists, 45(2):3–4, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1989:WRY

[Isa89m] John Isaacs. Washington report: You can’t keep a bad weapon down; Mother knows best …; but is anyone listening?; alas, no CAFE talks. Bulletin of the Atomic Scientists, 45(3):3–4, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Loc89b].

Isaacs:1990:BYB

[Isa90a] John Isaacs. Bulletin: 20-year battle on chemical weapons is over. Bulletin of the Atomic Scientists, 46(6):3–4, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kap90b].

Isaacs:1990:BLC

[Isa90b] John Isaacs. Bulletin: A little chemistry for the summit. Bulletin of the Atomic Scientists, 46(5):6, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BBA

[Isa90c] John Isaacs. Bulletin: Budget agony after Kuwait invasion. Bulletin of the Atomic Scientists, 46(9):3, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BCC

[Isa90d] John Isaacs. Bulletin: Calling Cousteau to account. Bulletin of the Atomic Scientists, 46(10):6, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BCD

[Isa90e] John Isaacs. Bulletin: Chernobyl data deleted. Bulletin of the Atomic Scientists, 46(10):6, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BCF

[Isa90f] John Isaacs. Bulletin: Conference fallout. Bulletin of the Atomic Scientists, 46(10):4, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BFB

[Isa90g] John Isaacs. Bulletin: Fast break on bioweapons. Bulletin of the Atomic Scientists, 46(6):5–6, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BHK

[Isa90h] John Isaacs. Bulletin: Hawks keep selling war. Bulletin of the Atomic Scientists, 46(2):5, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Whe90].

Isaacs:1990:BHS

[Isa90i] John Isaacs. Bulletin: Heady successes. Bulletin of the Atomic Scientists, 46(6):5, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BHC

[Isa90j] John Isaacs. Bulletin: House in from the cold. Bulletin of the Atomic Scientists, 46(8):4, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BHM

[Isa90k] John Isaacs. Bulletin: How the mighty have fallen. Bulletin of the Atomic Scientists, 46(10):6, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BIG

[Isa90l] John Isaacs. Bulletin: Iraq and the golden tongues. Bulletin of the Atomic Scientists, 46(8):3, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BMA

[Isa90m] John Isaacs. Bulletin: Madison Avenue tries selling peace. Bulletin of the Atomic Scientists, 46(2):4, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BNN

[Isa90n] John Isaacs. Bulletin: Nunn’s new thinking. Bulletin of the Atomic Scientists, 46(5):4–5, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BPW

[Isa90o] John Isaacs. Bulletin: Partners on world patrol. Bulletin of the Atomic Scientists, 46(2):6–7, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BSL

[Isa90p] John Isaacs. Bulletin: Seeing the light. Bulletin of the Atomic Scientists, 46(3):5–6, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BSSa

[Isa90q] John Isaacs. Bulletin: Senate stuck in 1980s. Bulletin of the Atomic Scientists, 46(8):3–4, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BSMb

[Isa90r] John Isaacs. Bulletin: Short memories. Bulletin of the Atomic Scientists, 46(9):4, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BSH

[Isa90s] John Isaacs. Bulletin: Soviet hearings on test ban. Bulletin of the Atomic Scientists, 46(10):7, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BSMa

[Isa90t] John Isaacs. Bulletin: START makes sense despite oversell. Bulletin of the Atomic Scientists, 46(5):3, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BSB

[Isa90u] John Isaacs. Bulletin: Sunscreens and bullion for the B-2. Bulletin of the Atomic Scientists, 46(3):4–5, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BTTa

[Isa90v] John Isaacs. Bulletin: Targeting targeting. Bulletin of the Atomic Scientists, 46(7):5–6, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BTTb

[Isa90w] John Isaacs. Bulletin: Testing treating slip through. Bulletin of the Atomic Scientists, 46(10):5, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:BGP

[Isa90x] John Isaacs. Bulletin: The greening of the Pentagon. Bulletin of the Atomic Scientists, 46(7):4–5, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rob91a].

Isaacs:1990:BSSb

[Isa90y] John Isaacs. Bulletin: The Saddam and Sam show. Bulletin of the Atomic Scientists, 46(10):3, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:RMB

[Isa90z] John Isaacs. Report: Military budget: This time, Congress speaks up. Bulletin of the Atomic Scientists, 46(4):6–7, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1990:WCC

[Isa90-27] John Isaacs. Washington: Clinging to the Cold War. Bulletin of the Atomic Scientists, 46(1):9–10, 44, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:ACW

[Isa91a] John Isaacs. Arms control: What a difference a year makes. Bulletin of the Atomic Scientists, 47(1):10–11, 45, January 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BCSa

[Isa91b] John Isaacs. Bulletin: China still favored. Bulletin of the Atomic Scientists, 47(7):7, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BCF

[Isa91c] John Isaacs. Bulletin: Congress in a funk. Bulletin of the Atomic Scientists, 47(3):4–5, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BCSb

[Isa91d] John Isaacs. Bulletin: Congress seizes Bush’s weapons initiative. Bulletin of the Atomic Scientists, 47(9):3–5, November 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BDA

[Isa91e] John Isaacs. Bulletin: Death to the Avenger. Bulletin of the Atomic Scientists, 47(2):5, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BDDa

[Isa91f] John Isaacs. Bulletin: Divided debate on a foregone conclusion. Bulletin of the Atomic Scientists, 47(2):3–5, March 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See date corrections [Kir91].

Isaacs:1991:BDDc

[Isa91g] John Isaacs. Bulletin: Divvying up the disarmament spoils. Bulletin of the Atomic Scientists, 47(10):4–5, December 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BDDb

[Isa91h] John Isaacs. Bulletin: Double duty for high-tech Gulf weapons. Bulletin of the Atomic Scientists, 47(3):3–4, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BFS

[Isa91i] John Isaacs. Bulletin: First shots in the counterrevolution. Bulletin of the Atomic Scientists, 47(4):3–4, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BNB

[Isa91j] John Isaacs. Bulletin: Northrop bungles B-2 PR package. Bulletin of the Atomic Scientists, 47(5):3, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BRR

[Isa91k] John Isaacs. Bulletin: Round and round on arms sales. Bulletin of the Atomic Scientists, 47(5):3–4, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BSS

[Isa91l] John Isaacs. Bulletin: Safety in secret. Bulletin of the Atomic Scientists, 47(2):5, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BSCa

[Isa91m] John Isaacs. Bulletin: SDI coming down to earth? Bulletin of the Atomic Scientists, 47(7):5–7, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BST

[Isa91n] John Isaacs. Bulletin: Shifting tide reshapes 1992 agenda. Bulletin of the Atomic Scientists, 47(10):5–6, December 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BSD

[Isa91o] John Isaacs. Bulletin: Snatching defeat from the jaws of victory. Bulletin of the Atomic Scientists, 47(6):3–4, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BSCb

[Isa91p] John Isaacs. Bulletin: START, CFE shoo-ins. Bulletin of the Atomic Scientists, 47(9):5–6, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:BRS

[Isa91q] John Isaacs. Bulletin: The remilitarization of Sam Nunn. Bulletin of the Atomic Scientists, 47(8):3–4, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1991:RCW

[Isa91r] John Isaacs. Report: Chemical weapons: We love them, we love them not. Bulletin of the Atomic Scientists, 47(6):9, 43, July/ August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BRM

[Isa92a] John Isaacs. Bulletin: A raging moderate. Bulletin of the Atomic Scientists, 48(7):3–4, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BAT

[Isa92b] John Isaacs. Bulletin: Aspin’s terrible swift sword. Bulletin of the Atomic Scientists, 48(7):4–6, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BBC

[Isa92c] John Isaacs. Bulletin: Bush, Clinton put future on hold. Bulletin of the Atomic Scientists, 48(9):3–5, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BBWa

[Isa92d] John Isaacs. Bulletin: Bush whacked by Wofford win. Bulletin of the Atomic Scientists, 48(1):3–4, January/February 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BCT

[Isa92e] John Isaacs. Bulletin: Congress tests the waters. Bulletin of the Atomic Scientists, 48(6):6–7, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BDD

[Isa92f] John Isaacs. Bulletin: Defense debate: Senate blinks. Bulletin of the Atomic Scientists, 48(5):6–7, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BDSb

[Isa92g] John Isaacs. Bulletin: Defense spending — workfare for the ’90s? Bulletin of the Atomic Scientists, 48(3):3–5, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BDSa

[Isa92h] John Isaacs. Bulletin: Doves seek Senate roost. Bulletin of the Atomic Scientists, 48(2):4–7, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BBWb

[Isa92i] John Isaacs. Bulletin: In brief: Warm afterglow?; na zdorovye!; the road to Plutoburg; recycling the gulag. Bulletin of the Atomic Scientists, 48(4):5, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BPG

[Isa92j] John Isaacs. Bulletin: Playing games with arms control. Bulletin of the Atomic Scientists, 48(5):4–5, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:BTC

[Isa92k] John Isaacs. Bulletin: Talk is cheap. Bulletin of the Atomic Scientists, 48(4):7, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:NPN

[Isa92l] John Isaacs. New president; new policy? Bulletin of the Atomic Scientists, 48(10):10, 52, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:PGB

[Isa92m] John Isaacs. Perspective: Give Bush some credit. Bulletin of the Atomic Scientists, 48(6):14–15, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1992:RNT

[Isa92n] John Isaacs. Report: Nuclear testing: The Senate that can say no. Bulletin of the Atomic Scientists, 48(8):6–8, October 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BBD

[Isa93a] John Isaacs. Bulletin: A billion-dollar bonanza. Bulletin of the Atomic Scientists, 49(6):3–5, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BAL

[Isa93b] John Isaacs. Bulletin: ACDA on the line. Bulletin of the Atomic Scientists, 49(5):6, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BCU

[Isa93c] John Isaacs. Bulletin: “Conversion” from the unconverted. Bulletin of the Atomic Scientists, 49(6):6–7, July/August 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BHO

[Isa93d] John Isaacs. Bulletin: House OK’s Russian aid. Bulletin of the Atomic Scientists, 49(7):3–4, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BNS

[Isa93e] John Isaacs. Bulletin: It’s a national security policy, stupid. Bulletin of the Atomic Scientists, 49(7):5–7, September 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BPCb

[Isa93f] John Isaacs. Bulletin: Pentagon clings to costly lifestyle. Bulletin of the Atomic Scientists, 49(3):3–5, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BSW

[Isa93g] John Isaacs. Bulletin: Star Wars reverts to maiden name. Bulletin of the Atomic Scientists, 49(6):8, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BBC

[Isa93h] John Isaacs. Bulletin: The base-closure waltz. Bulletin of the Atomic Scientists, 49(4):5–6, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BCM

[Isa93i] John Isaacs. Bulletin: The Clinton military budget. Bulletin of the Atomic Scientists, 49(4):7–8, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BPCa

[Isa93j] John Isaacs. Bulletin: The paper chase. Bulletin of the Atomic Scientists, 49(1):3–4, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:BTO

[Isa93k] John Isaacs. Bulletin: Three outrages. Bulletin of the Atomic Scientists, 49(5):6–7, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-13918630.html.

Isaacs:1993:BTW

[Isa93l] John Isaacs. Bulletin: Treading water. Bulletin of the Atomic Scientists, 49(5):3–5, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:PBB

[Isa93m] John Isaacs. Perspective: A bouquet for Bush. Bulletin of the Atomic Scientists, 49(2):10, 46–47, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dri93].

Isaacs:1993:PB

[Isa93n] John Isaacs. Perspective: Bottoms up. Bulletin of the Atomic Scientists, 49(9):12–13, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:RPJ

[Isa93o] John Isaacs. Report: Peacekeeping: Just put it on our tab. Bulletin of the Atomic Scientists, 49(8):7–9, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1993:SWP

[Isa93p] John Isaacs. Strengthening the world policeman. Bulletin of the Atomic Scientists, 49(10):14–15, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1994:BCL

[Isa94a] John Isaacs. Bulletin: Clinton’s line of defense. Bulletin of the Atomic Scientists, 50(3):7–8, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1994:BMR

[Isa94b] John Isaacs. Bulletin: Moratorium renewed. Bulletin of the Atomic Scientists, 50(3):6, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1994:BWD

[Isa94c] John Isaacs. Bulletin: What did they mean? Bulletin of the Atomic Scientists, 50(3):8, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1994:PCC

[Isa94d] John Isaacs. Perspective: A confederation of caution. Bulletin of the Atomic Scientists, 50(4):14–15, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HI94].

Isaacs:1994:PAC

[Isa94e] John Isaacs. Perspective: Another Cold War casualty. Bulletin of the Atomic Scientists, 50(2):14–15, 16, March/April 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1994:RBC

[Isa94f] John Isaacs. Report: Budget cuts: Not in my district. Bulletin of the Atomic Scientists, 50(5):13–15, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1994:RMB

[Isa94g] John Isaacs. Report: Military budget: Ducking the big debate. Bulletin of the Atomic Scientists, 50(1):10–11, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaryan:1994:AUT

[Isa94h] Drastamat Isaryan. Armenia: Until there are no more trees. Bulletin of the Atomic Scientists, 50(1):28–31, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaryan:1994:GOR

[Isa94i] Drastamat Isaryan. Gateway to oil riches. Bulletin of the Atomic Scientists, 50(1):30, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1995:RMSa

[Isa95a] John Isaacs. Report: Military spending: Hawks on the wing. Bulletin of the Atomic Scientists, 51(2):14–16, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1995:RMSb

[Isa95b] John Isaacs. Report: Military spending: Hawks take a hit. Bulletin of the Atomic Scientists, 51(3):12–13, May/June 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wee95].

Isaacs:1995:RMSc

[Isa95c] John Isaacs. Report: Military spending: Senate snubs Strom. Bulletin of the Atomic Scientists, 51(4):11–12, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1995:RSW

[Isa95d] John Isaacs. Report: Star Wars: Senate says yes, maybe. Bulletin of the Atomic Scientists, 51(6):11–12, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1995:RSA

[Isa95e] John Isaacs. Right says arms control wrong: Right-wing Republicans would erase 10 years of treaty-making. Bulletin of the Atomic Scientists, 51(5):18–19, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1995:RTA

[Isa95f] John Isaacs. Right turn ahead. Bulletin of the Atomic Scientists, 51(1):16–17, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1996:BH

[Isa96a] John Isaacs. A Byrd, not a hawk. Bulletin of the Atomic Scientists, 52(5):20–21, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1996:RTB

[Isa96b] John Isaacs. The representation-without-taxation blues. Bulletin of the Atomic Scientists, 52(1):14–15, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1996:RAC

[Isa96c] John Isaacs. Report: Arms control: The Senate’s fits and START. Bulletin of the Atomic Scientists, 52(2):13–14, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1996:RCW

[Isa96d] John Isaacs. Report: Chemical weapons convention: Right Wing targets treaty. Bulletin of the Atomic Scientists, 52(6):13–15, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kyl97].

Isaacs:1996:RMS

[Isa96e] John Isaacs. Report: Military spending: Debatable, but not debated. Bulletin of the Atomic Scientists, 52(4):12–13, July/ August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dor96].

Isaacs:1996:RSW

[Isa96f] John Isaacs. Report: Star Wars: Play it again, Bob. Bulletin of the Atomic Scientists, 52(3):9–10, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1997:RCK

[Isa97a] John Isaacs. Report: Congress: Kinda good, kinda bad. Bulletin of the Atomic Scientists, 53(1):12, 66, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1997:RCP

[Isa97b] John Isaacs. Report: Congress: Playing to win. Bulletin of the Atomic Scientists, 53(5):17–18, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1997:RCS

[Isa97c] John Isaacs. Report: Congress: Spending more, enjoying it less. Bulletin of the Atomic Scientists, 53(3):19–20, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1997:RNC

[Isa97d] John Isaacs. Report: NATO: Congress snoozes. Bulletin of the Atomic Scientists, 53(2):15–16, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1997:RST

[Isa97e] John Isaacs. Report: The Senate: Treaty tactics. Bulletin of the Atomic Scientists, 53(4):13–14, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1997:RWS

[Isa97f] John Isaacs. Report: Washington: Spinning to the right. Bulletin of the Atomic Scientists, 53(6):14–15, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1998:NSS

[Isa98a] John Isaacs. NATO: The Senate’s strange bedfellows. Bulletin of the Atomic Scientists, 54(1):32–34, January/February 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1998:RDS

[Isa98b] John Isaacs. Report: Defense spending: When the walls come tumbling down. Bulletin of the Atomic Scientists, 54(6):14–15, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1998:RMD

[Isa98c] John Isaacs. Report: Missile defense: Rumbles from Rumsfeld. Bulletin of the Atomic Scientists, 54(5):16–17, September/ October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1998:RSA

[Isa98d] John Isaacs. Report: Senate: Aiming at ABM. Bulletin of the Atomic Scientists, 54(2):14–15, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1998:RTB

[Isa98e] John Isaacs. Report: Test ban: Maybe this year, maybe not. Bulletin of the Atomic Scientists, 54(3):16–17, May/June 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1998:STB

[Isa98f] John Isaacs. Senate: Test ban prospects shaken. Bulletin of the Atomic Scientists, 54(4):40–41, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:1999:CHY

[Isa99a] John Isaacs. Congress: Having your cake. Bulletin of the Atomic Scientists, 55(4):19–20, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/4/19.

Isaacs:1999:DSF

[Isa99b] John Isaacs. Defense spending: F-22 in a tailspin. Bulletin of the Atomic Scientists, 55(5):14–15, September/October 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/5/14.

Isaacs:1999:FFA

[Isa99c] John Isaacs. Fortifying Fortress America. Bulletin of the Atomic Scientists, 55(2):24–25, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/2/24.

Isaacs:1999:MDB

[Isa99d] John Isaacs. Missile defense: It’s back. Bulletin of the Atomic Scientists, 55(3):26–28, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/3/26.

Isaacs:1999:PP

[Isa99e] John Isaacs. Pragmatism over principle. Bulletin of the Atomic Scientists, 55(1):16–17, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/16.

Isaacs:1999:TBF

[Isa99f] John Isaacs. Test ban fizzles. Bulletin of the Atomic Scientists, 55(6):21–22, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/6/21.

Isaacs:2000:DSO

[Isa00a] John Isaacs. Defense spending: Onward and upward. Bulletin of the Atomic Scientists, 56(3):20–21, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/3/20.

Isaacs:2000:LPP

[Isa00b] John Isaacs. Like peas in a pod. Bulletin of the Atomic Scientists, 56(4):22–23, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/4/22.

Isaacs:2000:NSN

[Isa00c] John Isaacs. National security: The next President’s full agenda. Bulletin of the Atomic Scientists, 56(6):16–17, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/6/16.

Isaacs:2000:RPT

[Isa00d] John Isaacs. Not ready for prime time. Bulletin of the Atomic Scientists, 56(5):20–21, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/20.

Isaacs:2000:PD

[Isa00e] John Isaacs. A political decision. Bulletin of the Atomic Scientists, 56(2):23–25, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/2/23.

Isaacs:2000:WGH

[Isa00f] John Isaacs. Where to go from here. Bulletin of the Atomic Scientists, 56(1):40–41, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/40.

Isaacs:2001:BIR

[Isa01a] John Isaacs. Bush II or Reagan III? Bulletin of the Atomic Scientists, 57(3):30–31, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/3/30.

Isaacs:2001:MDT

[Isa01b] John Isaacs. Missile defense: Two steps forward, two steps back. Bulletin of the Atomic Scientists, 57(4):20–21, July 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/4/20.

Isaacs:2001:NSO

[Isa01c] John Isaacs. National security: The ones to watch. Bulletin of the Atomic Scientists, 57(2):25–27, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/25.

Isaacs:2001:PA

[Isa01d] John Isaacs. Pebbles and all. Bulletin of the Atomic Scientists, 57(5):22–23, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/5/22.

Isaacs:2001:TCN

[Isa01e] John Isaacs. Terrorism: Costing out the new agenda. Bulletin of the Atomic Scientists, 57(6):17–19, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/6/17.

Isaacs:2002:BDD

[Isa02a] John Isaacs. Bomb deal a dud, part two. Bulletin of the Atomic Scientists, 58(1):26–27, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/26.

Isaacs:2002:CSO

[Isa02b] John Isaacs. Congress: The Senate’s one-vote difference. Bulletin of the Atomic Scientists, 58(5):17–18, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/5/17.

Isaacs:2002:DBM

[Isa02c] John Isaacs. Defense budget: More money than sense. Bulletin of the Atomic Scientists, 58(2):22–24, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/22.

Isaacs:2002:EHS

[Isa02d] John Isaacs. The elections: Hot seats. Bulletin of the Atomic Scientists, 58(6):15–16, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/6/15.

Isaacs:2002:NPR

[Isa02e] John Isaacs. Nuclear posture: Ready, aim, fire. Bulletin of the Atomic Scientists, 58(4):21, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/4/21.

Isaacs:2002:WTC

[Isa02f] John Isaacs. War on terror: Congress speaks up. Bulletin of the Atomic Scientists, 58(3):16–17, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/3/16.

Isaacs:2003:CCP

[Isa03a] John Isaacs. Climbing costs, plunging popularity. Bulletin of the Atomic Scientists, 59(6):17–18, November/December 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/6/17.

Isaacs:2003:CDS

[Isa03b] John Isaacs. Congress: Democrats speak up. Bulletin of the Atomic Scientists, 59(5):??, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/5/19.

Isaacs:2003:CGA

[Isa03c] John Isaacs. Congress goes AWOL. Bulletin of the Atomic Scientists, 59(3):20, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isaacs:2003:CVV

[Isa03d] John Isaacs. Congress: To vote or not to vote. Bulletin of the Atomic Scientists, 59(2):19–20, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/2/19.

Isaacs:2003:EWT

[Isa03e] John Isaacs. The elections: War on terror wins. Bulletin of the Atomic Scientists, 59(1):18–19, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/1/18.

Isaacs:2003:NWC

[Isa03f] John Isaacs. Nuclear weapons: Congress jumps off the banwagon. Bulletin of the Atomic Scientists, 59(4):20–21, July/ August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
4/20.

Isaacs:2003:WIC

[Isa03g] John Isaacs. War in Iraq: Congress goes AWOL. Bulletin of the Atomic Scientists, 59(3):??, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/3/20.

Isaacs:2004:BKF

[Isa04a] John Isaacs. Bush, Kerry, and foreign policy. Bulletin of the Atomic Scientists, 60(5):20–21, September/October 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/5/20.

Isaacs:2004:BMH

[Isa04b] John Isaacs. Bush meets his Waterloo. Bulletin of the Atomic Scientists, 60(1):26–27, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/1/26.

Isaacs:2004:DSN

[Isa04c] John Isaacs. Defense spending: No more blank checks. Bulletin of the Atomic Scientists, 60(3):18–19, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/18.

Isaacs:2004:EWD

[Isa04d] John Isaacs. Elections: What the dems must do. Bulletin of the Atomic Scientists, 60(2):??, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/16.

Isaacs:2004:EC

[Isa04e] John Isaacs. The elephant in the closet. Bulletin of the Atomic Scientists, 60(6):17–18, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/6/17.

Isaacs:2004:MDS

[Isa04f] John Isaacs. Missile defense: The saga continues. Bulletin of the Atomic Scientists, 60(4):??, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/17.

Isaacs:2005:RO

[Isa05a] John Isaacs. 2006: a race odyssey. Bulletin of the Atomic Scientists, 61(4):??, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/4/22.

Isaacs:2005:EV

[Isa05b] John Isaacs. Echoes of Vietnam. Bulletin of the Atomic Scientists, 61(2):24–25, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/2/24.

Isaacs:2005:IB

[Isa05c] John Isaacs. An indefensible budget. Bulletin of the Atomic Scientists, 61(3):21–23, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/21.

Isaacs:2005:OL

[Isa05d] John Isaacs. Opportunity lost. Bulletin of the Atomic Scientists, 61(1):21–22, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/1/21.

Isaacson:2007:EHL

[Isa07] Walter Isaacson. Einstein: his life and universe. Simon and Schuster, New York, NY, USA, 2007. ISBN 0-7432-6473-8. xxii + 675 + 16 pp. LCCN QC16.E5 I76 2007. URL http://www.

loc.gov/catdir/enhancements/fy0708/2006051264-d.html.

Isenberg:1992:BRP

[Ise92] David Isenberg. Book review: Post-game analysis: Closing Pandora’s Box: Arms Races, Arms Control and the History of the Cold War, by Patrick Glynn. Bulletin of the Atomic Scientists, 48(9):41, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isenberg:1997:LDD

[Ise97] David Isenberg. Letter: Done deal. Bulletin of the Atomic Scientists, 53(4):59, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Khr97].

Isenberg:2001:GP

[Ise01] David Isenberg. “Guinea pigs”. Bulletin of the Atomic Scientists, 57(3):72–73, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/3/72.

Istock:1967:FUD

[Ist67] Conrad Istock. Forum: U.S. deaths in Vietnam: A definition of escalation. Bulletin of the Atomic Scientists, 23(6):62–63, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Istock:1998:PBM

[Ist98] Conrad A. Istock. Perspective: Bad medicine. Bulletin of the Atomic Scientists, 54(6):21–23, November/December 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Itonaga:2014:CCS

[Ito14] Koji Itonaga. Contamination and community support in the aftermath of the Fukushima disaster. Bulletin of the Atomic Scientists, 70(4):49–56, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ivanova:2002:BGM

[Iva02] Youliana Ivanova. Bulgaria: Goodbye missiles, hello NATO. Bulletin of the Atomic Scientists, 58(5):19–21, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/5/19.

Ivers:2016:ASL

[Ive16] Andrew Ivers. Adam Segal: Life in the hacked world order. Bulletin of the Atomic Scientists, 72(5):267–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ivy:1964:LEM

[Ivy64] A. C. Ivy. Letter to the Editor: More about Krebiozen. Bulletin of the Atomic Scientists, 20(7):28, September 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar64g].

Isard:1949:APE

[IW49] Walter Isard and Vincent Whitney. Atomic power and economic development. Bulletin of the Atomic Scientists, 5(3):73–79, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Isard:1952:APR

[IW52] Walter Isard and Vincent Whitney. Atomic power and regional development. Bulletin of the Atomic Scientists, 8(4):119–124, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Iwamatsu:1982:NPW

[Iwa82] Shigetoshi Iwamatsu. From Nagasaki: A perspective on the war crimes. Bulletin of the Atomic Scientists, 38(2):29–32, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jac82a].

Jack:1955:AAC

[Jac55] Homer A. Jack. The Asian–African conference. Bulletin of the Atomic Scientists, 11(6):221–222, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jackson:1956:ITB

[Jac56] Henry M. Jackson. The increasing threat of ballistic missiles. Bulletin of the Atomic Scientists, 12(3):90–92, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1958:TCB

[Jac58] Homer A. Jack. The Tokyo Conference and beyond. Bulletin of the Atomic Scientists, 14(2):91–92, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1961:BC

[Jac61] Homer A. Jack. The Belgrade Conference. Bulletin of the Atomic Scientists, 17(9):390–392, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1962:MPC

[Jac62a] Homer A. Jack. Moscow Peace Congress: Criticism allowed. Bulletin of the Atomic Scientists, 18(8):42–43, October 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1962:UDR

[Jac62b] Homer A. Jack. U.N. disarmament resolutions. Bulletin of the Atomic Scientists, 18(4):17, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1963:DUQ

[Jac63a] Homer A. Jack. Disarmament at the U.N.: The quiet assembly. Bulletin of the Atomic Scientists, 19(2):39–40, 42–44, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1963:OCO

[Jac63b] Homer A. Jack. Oxford conference: Organizing the non-aligned. Bulletin of the Atomic Scientists, 19(6):38–40, June 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1964:DUD

[Jac64a] Homer A. Jack. Disarmament at the U.N.: Discussions in detente. Bulletin of the Atomic Scientists, 20(2):33–36, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1964:CNC

[Jac64b] Homer A. Jack. From Cairo: The nonaligned confer. Bulletin of the Atomic Scientists, 20(10):34–35, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1964:SC

[Jac64c] Homer A. Jack. Seventeen continue. Bulletin of the Atomic Scientists, 20(6):43–45, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1965:UWD

[Jac65] Homer A. Jack. UNCD: For a world disarmament conference. Bulletin of the Atomic Scientists, 21(7):39–40, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1966:TWD

[Jac66] Homer A. Jack. Toward a world disarmament conference. Bulletin of the Atomic Scientists, 22(4):47–48, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1967:REG

[Jac67] Homer A. Jack. Reports: ENDC at the General Assembly. Bulletin of the Atomic Scientists, 23(2):30–33, February 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacob:1968:RMG

[Jac68] François Jacob. Reports: Microbiology and genetics. Bulletin of the Atomic Scientists, 24(7):41–48, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jackson:1969:CWS

[Jac69] Henry M. Jackson. Czechoslovakia and Western security. Bulletin of the Atomic Scientists, 25(1):36–39, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1971:GAV

[Jac71a] Sally Jacobsen. Getting aboard Viking: No room on the Mars lander. Bulletin of the Atomic Scientists, 27(10):49–50, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1971:NFC

[Jac71b] Sally Jacobsen. Notes on the Fourth Conference on Plasma Physics and Nuclear Fusion Research at Madison, Wis., June 17–23, 1971. Bulletin of the Atomic Scientists, 27(8):49, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1972:BSR

[Jac72a] Sally Jacobsen. A Bulletin special report: Turning up the gas: AEC prepares another nuclear gas stimulation shot. Bulletin of the Atomic Scientists, 28(5):35–38, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1972:SRW

[Jac72b] Sally Jacobsen. A special report: What happened at Stockholm. II. A call to environmental order. Bulletin of the Atomic Scientists, 28(7):21–25, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1973:WDC

[Jac73a] Homer A. Jack. A world disarmament conference? Bulletin of the Atomic Scientists, 29(2):33–35, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1973:BSRa

[Jac73b] Sally Jacobsen. A Bulletin special report: Land use dispute in Illinois: Nuclear power vs. crops. Bulletin of the Atomic Scientists, 29(1):43–45, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1973:BSRb

[Jac73c] Sally Jacobsen. Bulletin special report: The Great Montana Coal Rush. Bulletin of the Atomic Scientists, 29(4):37–42, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1973:JVA

[Jac73d] Sally Jacobsen. James A. Van Allen: Research review: The trip to Jupiter. Bulletin of the Atomic Scientists, 29(10):52–56, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1973:RRH

[Jac73e] Sally Jacobsen. Research review: Hannes Alfven on space exploration. Bulletin of the Atomic Scientists, 29(2):30–32, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1973:RRM

[Jac73f] Sally Jacobsen. Research review: Maurice F. Strong: Stockholm — a year later. Bulletin of the Atomic Scientists, 29(6):35–40, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1973:WOR

[Jac73g] Sally Jacobsen. Walter Orr Roberts on the atmosphere, global pollution and weather modification. Bulletin of the Atomic Scientists, 29(3):42–46, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1974:AWH

[Jac74a] Sally Jacobsen. Addendum: The White House on cooling towers. Bulletin of the Atomic Scientists, 30(1):45, January 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1974:MMA

[Jac74b] Sally Jacobsen. Minutes to midnight: The arms race: Bigger Pentagon budgets, more Soviet ICBM shots; no on the genocide pact; an ‘animal fear of exposures”; the energy crisis: A body count; the Cultural Revolution, phase 2; a Vatican switch on birth control?; nuclear moratorium gains a convert; AEC says reactors are safe; the lion’s share for LMFBR; sewage funding backed up; full speed ahead on the pipeline. Bulletin of the Atomic Scientists, 30(4):4–6, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1974:RSA

[Jac74c] Sally Jacobsen. Reflections of a science advisor: An interview with Edward E. David, Jr. Bulletin of the Atomic Scientists, 30 (2):27–28, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1974:SRA

[Jac74d] Sally Jacobsen. Special report: Anti-pollution backlash in Illinois: Can a tough protection program survive? Bulletin of the Atomic Scientists, 30(1):39–44, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1974:TCJ

[Jac74e] Sally Jacobsen. A time for change for the Joint Committee. Bulletin of the Atomic Scientists, 30(5):9–11, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jack:1977:DS

[Jac77] Homer A. Jack. The disarmament scoreboard. Bulletin of the Atomic Scientists, 33(3):54–58, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jackson:1979:PAC

[Jac79] William D. Jackson. Policy assessment at the crossroads: the Soviets and SALT. Bulletin of the Atomic Scientists, 35(4):10–14, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1980:CCA

[Jac80a] C. G. Jacobsen. Commentary from Canada: Afghanistan: ice and fire. Bulletin of the Atomic Scientists, 36(3):5–6, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1980:BRB

[Jac80b] Carl Jacobsen. Books review: Asia’s Nuclear Future, edited by William Overholt; The Military Equation in Northeast Asia, by Stuart E. Johnson and Joseph A. Yager; Strategy and Security in Northeast Asia, by Richard B. Foster, James E. Dornan, Jr., and William M. Carpenter. Bulletin of the Atomic Scientists, 36 (5):51–53, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jackson:1982:CAO

[Jac82a] Terry W. Jackson. Commentary: Another opinion on the war crimes. Bulletin of the Atomic Scientists, 38(4):54, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Iwa82].

Jackson:1982:CWD

[Jac82b] William D. Jackson. Cold War demonology. Bulletin of the Atomic Scientists, 38(8):52–54, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jackson:1982:RUS

[Jac82c] William E. Jackson, Jr. Reagan’s unsavory SALT. Bulletin of the Atomic Scientists, 38(7):15–16, August/September 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1982:CAD

[Jac82d] Carl G. Jacobsen. Commentary: Allies divided. Bulletin of the Atomic Scientists, 38(1):46–48, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1983:CPV

[Jac83] Carl G. Jacobsen. Commentary: The Puritan vision of arms control. Bulletin of the Atomic Scientists, 39(8):60–61, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1992:BRL

[Jac92] Carl G. Jacobsen. Book review: The last struggle: The August Coup: The Truth and the Lessons, by Mikhail Gorbachev. Bulletin of the Atomic Scientists, 48(1):41–42, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacob:2015:SLR

[Jac15] Klaus H. Jacob. Sea level rise, storm risk, denial, and the future of coastal cities. Bulletin of the Atomic Scientists, 71(5): 40–50, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/40.full.pdf+html.

Jaffey:1951:CDNa

[Jaf51a] Arthur H. Jaffey. Civil defense news. Bulletin of the Atomic Scientists, 7(6):192, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jaffey:1951:CDNb

[Jaf51b] Arthur H. Jaffey. Civil defense news. Bulletin of the Atomic Scientists, 7(7–8):234–235, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jagger:1995:LTD

[Jag95] John Jagger. Letter: They’re not dumps. Bulletin of the Atomic Scientists, 51(1):3, 65, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano94l] and comment [Blo95].

Jagger:1998:LSB

[Jag98] John Jagger. Letter: Sierra Blanca is safe. Bulletin of the Atomic Scientists, 54(4):67–68, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet98].

Jaipal:1973:CGE

[Jai73] Rikhi Jaipal. Communications: Genetic engineering. Bulletin of the Atomic Scientists, 29(5):2, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

James:1958:SEE

[Jam58] Sir Eric James. Science and education: An English view. Bulletin of the Atomic Scientists, 14(9):359–363, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jamgotch:1986:SCO

[Jam86] Nish Jamgotch, Jr. Superpower cooperation often overlooked. Bulletin of the Atomic Scientists, 42(2):37–39, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janis:1950:PPB

[Jan50] Irving L. Janis. Psychological problems of A-bomb defense. Bulletin of the Atomic Scientists, 6(8–9):256–262, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Janson:1959:ABW

[Jan59] H. W. Janson. After Betsy, what? Bulletin of the Atomic Scientists, 15(2):68–71, 93, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janowitz:1964:MPD

[Jan64] Morris Janowitz. The military in the political development of new nations. Bulletin of the Atomic Scientists, 20(8):6–10, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janouch:1973:CFS

[Jan73] F. Janouch. Communications: A “Free” slave. Bulletin of the Atomic Scientists, 29(7):55–56, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ezr72] and comment [Ano74b].

Janifer:1976:LLB

[Jan76a] Laurence M. Janifer. Letter: Lost balance-wheel. Bulletin of the Atomic Scientists, 32(8):2, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janouch:1976:LN

[Jan76b] Franišek Janouch. Letter: Nepřípustné. Bulletin of the Atomic Scientists, 32(5):4, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janouch:1976:SUS

[Jan76c] František Janouch. Science under siege in Czechoslovakia. Bulletin of the Atomic Scientists, 32(4):6–12, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [HJ76].

Janouch:1978:ISA

[Jan78] František Janouch. [Introduction to Sakharov article]. Bulletin of the Atomic Scientists, 34(6):12, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janouch:1980:CSS

[Jan80] F. Janouch. Commentary: Support for Sakharov. Bulletin of the Atomic Scientists, 36(9):64, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janeway:1981:CJL

[Jan81a] Eliot Janeway. Commentary: That Janeway letter. Bulletin of the Atomic Scientists, 37(8):63–64, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janowitz:1981:CMA

[Jan81b] Morris Janowitz. Commentary: Making the all-volunteer military work? Bulletin of the Atomic Scientists, 37(2):7–8, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janberg:2015:PRB

[Jan15] Klaus Janberg. Plutonium reprocessing, breeder reactors, and decades of debate : a German response. Bulletin of the Atomic Scientists, 71(4):10–13, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jasanoff:1981:BRB

[Jas81] Sheila Jasanoff. Book reviews: Pollution, Politics, and International Law — Tankers at Sea, by R. Michael M’Gonigle and Mark W. Zacher; Seabrook and the Nuclear Regulatory Commission — The Licensin of a Nuclear Power Plant, by Donald W. Stever, Jr.; Compliance and Public Authority — A Theory with International Applications, by Oran R. Young. Bulletin of the Atomic Scientists, 37(7):44–46, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jasper:1991:LNP

[Jas91] James M. Jasper. Letter: Nuclear power politics. Bulletin of the Atomic Scientists, 47(10):43, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SB91].

Jassby:2024:FSQ

[Jas24] Daniel L. Jassby. The fuel supply quandary of fusion power reactors. Bulletin of the Atomic Scientists, 80(6):377–384, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2418708.

Jauchem:1990:BME

[Jau90a] J. R. Jauchem. Brodeur misinformed on ELFs. Bulletin of the Atomic Scientists, 46(9):46, November 1990. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Har91a].

Jauchem:1990:LBM

[Jau90b] James R. Jauchem. Letter: Brodeur misinformed on ELFs. Bulletin of the Atomic Scientists, 46(9):46, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha90].

Javits:1968:DND

[Jav68] Jacob Javits. Dynamics of a national decision. Bulletin of the Atomic Scientists, 24(2):18–19, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Joliot-Curie:1950:PTE

[JC50] Frédéric Joliot-Curie. A proposal toward the elimination of the atomic danger. Bulletin of the Atomic Scientists, 6(6):166–167, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Joliot-Curie:1954:LEM

[JC54] I. Joliot-Curie. Letter to the Editor: On my rejection for membership in the American Chemical Society. Bulletin of the Atomic Scientists, 10(8):335, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jasinski:2002:RTT

[JCF02] Michael Jasinski, Christina Chuen, and Charles D. Ferguson. Russia: of truth and testing. Bulletin of the Atomic Scientists, 58(5):60–65, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/5/60.

Joliot-Curie:1950:FAS

[JCKG^{+}50] Frédéric Joliot-Curie, L. Kowarski, J. Gueron, B. Goldschmidt, J. Stohr, and M. Surdin. French atomic scientists report on their work in 1949. Bulletin of the Atomic Scientists, 6(10):299–302, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jeffries:1994:LAT

[Jef94] Marjorie S. Jeffries. Letter: Authentic thanks. Bulletin of the Atomic Scientists, 50(3):61, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zag94e].

Jehle:1960:LEB

[Jeh60a] Herbert Jehle. Letter to the Editor: Ban the bomb. Bulletin of the Atomic Scientists, 16(3):80, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jehle:1960:LEN

[Jeh60b] Herbert Jehle. Letter to the Editor: Nuclear test ban and detection of underground tests. Bulletin of the Atomic Scientists, 16(9):344, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jennison:1963:ABA

[Jen63] Peter S. Jennison. American books abroad. Bulletin of the Atomic Scientists, 19(10):31–33, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jensen:1970:JTT

[Jen70] Arthur R. Jensen. The Jensen thesis: Three comments: 1. Race and the genetics of intelligence: A reply to Lewontin. Bulletin of the Atomic Scientists, 26(5):17–23, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bha70].

Jen:1974:MSP

[Jen74] C. K. Jen. Mao’s “Serve the people” ethic. Bulletin of the Atomic Scientists, 30(3):15–25, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jerome:1977:FD

[Jer77] D. Jerome. Futility of deterrence. Bulletin of the Atomic Scientists, 33(8):8, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jervis:1985:MBP

[Jer85] Robert Jervis. MAD is the best possible deterrence. Bulletin of the Atomic Scientists, 41(3):43–45, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jervis:1987:BRB

[Jer87] Robert Jervis. Book review: The Rise of American Air Power, by Michael S. Sherry. Bulletin of the Atomic Scientists, 43(8): 51–52, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burbach:2019:OST

[JFB19] Joan Johnson-Freese and David Burbach. The Outer Space Treaty and the weaponization of space. Bulletin of the Atomic Scientists, 75(4):137–141, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2019.
1628458.

Johnson-Freese:2020:RSC

[JFS20] Joan Johnson-Freese and Nikola Schmidt. Reaching for the stars: The case for cooperative governance of directed energy technologies. Bulletin of the Atomic Scientists, 76(3):150–155, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1751972.

Jagger:1996:LWA

[JG96] John Jagger and Mark Goodman. Letters: We’re all radioactive. Bulletin of the Atomic Scientists, 52(2):3, 60, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gor96f].

Jeremias:2016:CEH

[JH16] Gunnar Jeremias and Mirko Himmel. Can everyone help verify the bioweapons convention? Perhaps, via open source monitoring. Bulletin of the Atomic Scientists, 72(6):412–417, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1974:MND

[JL74] David Johnson and Gene La Rocque. The mythology of national defense: How U.S. strategists justify the momentum of new weaponry. Bulletin of the Atomic Scientists, 30(7):21–26, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cae74].

Jones:1987:CAB

[JL87] R. V. Jones and J. M. Lewis. Churchill’s anthrax bombs: a debate. Bulletin of the Atomic Scientists, 43(9):42–, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber87c].

Jassby:1983:CNP

[JM83] D. L. Jassby and Philip Morrison. Commentary: On ‘The Nuclear Plateau’. Bulletin of the Atomic Scientists, 39(3):62, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mor82b].

Johnson:1985:LJW

[JM85] James Turner Johnson and Richard Miller. Letter: Just war. Bulletin of the Atomic Scientists, 41(9):56–57, October 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mil85b].

Johnson:1951:BBF

[Joh51] Walter Johnson. Books: The Forrestal Diaries, edited by Walter Millis with the collaboration of E. S. Duffield. Bulletin of the Atomic Scientists, 7(11):347–348, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1958:GSM

[Joh58] Frances R. Johnson. [Government support of mathematics]. Bulletin of the Atomic Scientists, 14(2):90, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from the author’s article Gresham College: Precursor of the Royal Society, Journal of the History of Ideas, 1(4) October 1940, 413–438.

Johnson:1959:NFP

[Joh59] Jesse C. Johnson. Nuclear fuel: Prospects and production. Bulletin of the Atomic Scientists, 15(6):245–248, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1960:NES

[Joh60a] Gerald W. Johnson. Nuclear explosions in science and technology. Bulletin of the Atomic Scientists, 16(5):155–161, May 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1960:SPI

[Joh60b] Gerald W. Johnson. The Soviet program for industrial application of explosions. Bulletin of the Atomic Scientists, 16(9):366–369, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1964:LES

[Joh64a] D. Gale Johnson. Letter to the Editor: Soviet agriculture. Bulletin of the Atomic Scientists, 20(6):35–36, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1964:SA

[Joh64b] D. Gale Johnson. Soviet agriculture. Bulletin of the Atomic Scientists, 20(1):8–12, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1964:CW

[Joh64c] Lyndon B. Johnson. [The cold war]. Bulletin of the Atomic Scientists, 20(7):7, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

John:1965:BHI

[Joh65a] E. Roy John. The brain and how it changes. Bulletin of the Atomic Scientists, 21(9):12–14, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Based on a paper The Nuclear Scientific Era: The Child and His Education, presented at a conference at the Walden School, New York City, April 24–26, 1965 [see [Ano66g]].

Johnson:1965:PBR

[Joh65b] Harry G. Johnson. Paying for basic research: Some economic issues. Bulletin of the Atomic Scientists, 21(10):12–16, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1965:FTC

[Joh65c] Walter Johnson. Have faith in the Twentieth Century. Bulletin of the Atomic Scientists, 21(1):10–15, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnston:1966:TI

[Joh66] J. Richard Johnston. The teach-ins. Bulletin of the Atomic Scientists, 22(2):30–31, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1967:BM

[Joh67a] Lyndon B. Johnson. From the budget message. Bulletin of the Atomic Scientists, 23(6):9, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1967:SUM

[Joh67b] Lyndon B. Johnson. From the State of the Union message. Bulletin of the Atomic Scientists, 23(6):8, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1968:NTJ

[Joh68] Lyndon B. Johnson. On the Nonproliferation Treaty, July 1, 1968. Bulletin of the Atomic Scientists, 24(7):18, September 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1970:PC

[Joh70] Gerald W. Johnson. Plowshare at the crossroads. Bulletin of the Atomic Scientists, 26(6):83–91, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1979:MAS

[Joh79] Giff Johnson. Micronesia: America’s ‘strategic’ trust. Bulletin of the Atomic Scientists, 35(2):10–15, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1980:CPN

[Joh80a] Giff Johnson. Commentary: Pacific nuclear waste storage. Bulletin of the Atomic Scientists, 36(9):62–63, November 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1980:PL

[Joh80b] Giff Johnson. Paradise lost. Bulletin of the Atomic Scientists, 36 (10):24–29, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [MJ81, Alc81].

Johnson:1981:FRB

[Joh81a] Baylor L. Johnson. Film review: Unforgettable Fire: Pictures Drawn by Atomic Bomb Survivors. Bulletin of the Atomic Scientists, 37(7):43–44, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1981:CLC

[Joh81b] Brian Johnson. Commentary from London: Collision course: conservation and economic survival. Bulletin of the Atomic Scientists, 37(4):11–12, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1983:STO

[Joh83a] Christopher Johnson. Students and teachers opposed to nuclear war. Bulletin of the Atomic Scientists, 39(9):61, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1983:SUA

[Joh83b] D. Gale Johnson. The Soviet Union: Agriculture — management and performance. Bulletin of the Atomic Scientists, 39(2):16– 22, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1983:CDD

[Joh83c] David T. Johnson. The case for delayed deployment. Bulletin of the Atomic Scientists, 39(10):30, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1986:SAN

[Joh86] Thomas B. Johansson. Sweden’s abortive nuclear weapons project. Bulletin of the Atomic Scientists, 42(3):31–34, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho86b].

Johnson:1989:TWP

[Joh89] Robert H. Johnson. Third world: Playing to the home crowd. Bulletin of the Atomic Scientists, 45(1):24–28, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1991:NWO

[Joh91] Leonard V. Johnson. New world order: Time for common security. Bulletin of the Atomic Scientists, 47(5):28, June 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1993:BSR

[Joh93] Sarah Johansson. Bulletin: Spatial relations. Bulletin of the Atomic Scientists, 49(9):3–4, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1994:AFC

[Joh94a] Sarah Johansson. Air Force carries Milstar around its neck. Bulletin of the Atomic Scientists, 50(2):7–8, March/April 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1994:BBB

[Joh94b] Sarah Johansson. Bulletin: Big Brother’s new hearing aids. Bulletin of the Atomic Scientists, 50(3):4–6, May/June 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1994:BCT

[Joh94c] Sarah Johansson. Bulletin: Congress turns down the radio. Bulletin of the Atomic Scientists, 50(1):6–7, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1994:BRR

[Joh94d] Sarah Johansson. Bulletin: Russian reform: for women, it’s not working. Bulletin of the Atomic Scientists, 50(4):5–6, July/ August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1994:BSR

[Joh94e] Sarah Johansson. Bulletin: Scrapping the rules. Bulletin of the Atomic Scientists, 50(2):6–7, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johansson:1994:BWF

[Joh94f] Sarah Johansson. Bulletin: We fry harder. Bulletin of the Atomic Scientists, 50(3):10–11, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1995:RTB

[Joh95] Rebecca Johnson. Report: Test ban treaty: Rearranging deck chairs on the Titanic. Bulletin of the Atomic Scientists, 51(5): 11–12, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1996:GBS

[Joh96a] Chalmers Johnson. Go-banken-sama, go home! Bulletin of the Atomic Scientists, 52(4):22–25, 27–29, July/August 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1996:CTB

[Joh96b] Rebecca Johnson. The in-comprehensive test ban. Bulletin of the Atomic Scientists, 52(6):30–35, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1996:LVG

[Joh96c] Rebecca Johnson. … and the view from Geneva. Bulletin of the Atomic Scientists, 52(1):44–45, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1997:BRD

[Joh97a] Amanda Johnson. Book review: Dark secrets: Hidden Horrors: Japanese War Crimes in World War II, by Yuki Tanaka. Bulletin of the Atomic Scientists, 53(1):62–63, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1997:CW

[Joh97b] Chalmers Johnson. The Chinese way. Bulletin of the Atomic Scientists, 53(1):20–25, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1997:GOL

[Joh97c] Rebecca Johnson. Guest opinion: Little Orphan Fissban. Bulletin of the Atomic Scientists, 53(3):4, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Johnson:1999:SNC

[Joh99a] Chalmers Johnson. In search of a new Cold war. Bulletin of the Atomic Scientists, 55(5):44–51, September/October 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/5/44.

Johnson:1999:TTN

[Joh99b] Rebecca Johnson. Troubled treaties: Is the NPT tottering? Bulletin of the Atomic Scientists, 55(2):16–18, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/2/16.

Johnson:2000:NRD

[Joh00a] Rebecca Johnson. The NPT review: Disaster averted. Bulletin of the Atomic Scientists, 56(4):52–57, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/4/52.

Johnson:2000:WT

[Joh00b] Rebecca Johnson. A world of trouble. Bulletin of the Atomic Scientists, 56(1):38–39, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/38.

Johnson:2001:MDD

[Joh01] Rebecca Johnson. Missile defense: Downing Street says yes; Britons, no. Bulletin of the Atomic Scientists, 57(6):28–30, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
6/28.

Johnson:2003:GM

[Joh03] Loch K. Johnson. The G-Man. Bulletin of the Atomic Scientists, 59(1):67–68, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/67.

Johnson:2004:BPA

[Joh04] Rebecca Johnson. Britain: Pro-American no more? Bulletin of the Atomic Scientists, 60(2):18–21, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/18.

Johnson:2005:ACA

[Joh05] Rebecca Johnson. Arms control: All talk, no action. Bulletin of the Atomic Scientists, 61(4):19–20, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/4/19.

Johnson:2010:ANR

[Joh10] Rebecca Johnson. Assessing the 2010 NPT review conference. Bulletin of the Atomic Scientists, 66(4):1–10, July/August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/4/1.

Johnson:2022:SFF

[Joh22] Kaitlyn Johnson. Space force: Fact or fiction? Bulletin of the Atomic Scientists, 78(1):22–25, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2021.
2014232.

JC:1954:CRA

[Joi54] Joint Committee. Calculated risks in the AEC. Bulletin of the Atomic Scientists, 10(7):263, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Joliot:1946:FAE

[Jol46] Frédéric Joliot. France and atomic energy. Bulletin of the Atomic Scientists, 1(10):5, 7, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jonas:1959:AES

[Jon59] Anne M. Jonas. Atomic energy in Soviet bloc nations. Bulletin of the Atomic Scientists, 15(9):379–383, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1963:LER

[Jon63a] Arthur W. Jones. Letter to the Editor: Review articles. Bulletin of the Atomic Scientists, 19(1):32, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1963:LIM

[Jon63b] Harry W. Jones. Law and the idea of mankind. Bulletin of the Atomic Scientists, 19(1):6–12, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1965:LES

[Jon65] Galen E. Jones. The living economy of the sea. Bulletin of the Atomic Scientists, 21(3):13–17, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1970:DUA

[Jon70] Thomas O. Jones. Developing the U.S. Antarctic research program. Bulletin of the Atomic Scientists, 26(10):81–84, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Jones:1971:CEC

[Jon71] B. K. Jones. Communications: Electric car. Bulletin of the Atomic Scientists, 27(2):2, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Net70].

Jonas:1984:IRS

[Jon84] Hans Jonas. The imperative of responsibility: in search of an ethics for the technological age. University of Chicago Press, Chicago, IL, USA and London, UK, 1984. ISBN 0-226-40596-6. xii + 255 pp. LCCN BJ1453 .J6613 1984. Translated by Hans Jonas and with the collaboration of David Herr.

Jones:1987:BRB

[Jon87] David R. Jones. Book review: The Myth of Soviet Military Supremacy, by Tom Gervasi. Bulletin of the Atomic Scientists, 43 (5):51–52, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1990:LSM

[Jon90a] Jeff M. Jones. Letter: Slash military budget now, not in 10 years. Bulletin of the Atomic Scientists, 46(5):42, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1990:BRR

[Jon90b] Rodney W. Jones. Book review: Regional conflict in the Middle East: Nuclearisation of the Middle East, by Anoushiravan Ehteshami. Bulletin of the Atomic Scientists, 46(1):43–44, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1991:BHW

[Jon91a] Michael Jones. Bulletin: Hawaii wants no star wars STARS. Bulletin of the Atomic Scientists, 47(1):3, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1991:LST

[Jon91b] T. L. Jones. Letter: Superpowers too. Bulletin of the Atomic Scientists, 47(3):47, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi90b].

Jones:1993:RSW

[Jon93a] Michael Jones. Report: Star Wars: STARS no star on Kauai. Bulletin of the Atomic Scientists, 49(3):11–13, April 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1993:LBI

[Jon93b] T. L. Jones. Letter: Beef up the IAEA. Bulletin of the Atomic Scientists, 49(7):55, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha93b].

Jones:1997:LNP

[Jon97] Rodney W. Jones. Letter: No principle here. Bulletin of the Atomic Scientists, 53(1):3, 67, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo96i].

Jones:2013:CDF

[Jon13] Nate Jones. Countdown to declassification: Finding answers to a 1983 nuclear war scare. Bulletin of the Atomic Scientists, 69 (6):47–57, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/6/47.

Jamieson:2015:SPC

[JOO15] Dale Jamieson, Naomi Oreskes, and Michael Oppenheimer. Science and policy: Crossing the boundary. Bulletin of the Atomic Scientists, 71(1):53–58, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jordan:1963:FIR

[Jor63] Nehemiah Jordan. Forum: International relations and the psychologist. Bulletin of the Atomic Scientists, 19(9):29–33, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply: [SJ64].

Jordan:1967:LEM

[Jor67] Nehemiah Jordan. Letter to the Editor: Mind, brain, human values. Bulletin of the Atomic Scientists, 23(1):33–34, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spe66, GJ67].

Joravsky:1987:BRB

[Jor87] David Joravsky. Book review: The Soviet Union: The Incomplete Superpower, by Paul Dibb. Bulletin of the Atomic Scientists, 43(2):53–54, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jorant:2011:IFE

[Jor11] Caroline Jorant. The implications of Fukushima: The European perspective. Bulletin of the Atomic Scientists, 67(4):14–17, July/ August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
4/14.

Jordaan:2018:RPS

[Jor18] Sarah M. Jordaan. Resilience for power systems amid a changing climate. Bulletin of the Atomic Scientists, 74(2):95–101, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1436810.

Joshi:1961:DAI

[Jos61] Madhu Joshi. Dead or alive? International Atomic Energy Agency. Bulletin of the Atomic Scientists, 17(3):95–98, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Joseph:1981:CDS

[Jos81] Peter G. Joseph. Commentary: Doctors speak up. Bulletin of the Atomic Scientists, 37(3):17, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Josephson:1987:EYS

[Jos87] Paul R. Josephson. Early years of Soviet nuclear physics. Bulletin of the Atomic Scientists, 43(10):36–39, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Josephson:2002:MDG

[Jos02] Paul Josephson. Minatom: Dreams of glory. Bulletin of the Atomic Scientists, 58(5):40–47, September/October 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/5/40.

Joule:1952:IAW

[Jou52] James Prescott Joule. “The improvement of the art of war”. Bulletin of the Atomic Scientists, 8(6):200, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jacobsen:1982:FL

[JP82] Carl G. Jacobsen and Frank Press. Furthermore …. Bulletin of the Atomic Scientists, 38(3):9, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Prz82].

Jacoby:2013:NEU

[JP13] Henry D. Jacoby and Sergey Paltsev. Nuclear exit, the US energy mix, and carbon dioxide emissions. Bulletin of the Atomic Scientists, 69(2):34–43, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/69/2/34.

Jasani:1990:PSS

[JPR90] Bhupendra Jasani, Gwyn Prins, and Martin Rees. Perspective: Share satellite surveillance. Bulletin of the Atomic Scientists, 46(2):15–16, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jungk:1962:LEH

[JR62] Robert Jungk and John J. Robinson. Letter to the Editor: Hiroshima. Bulletin of the Atomic Scientists, 18(9):29–30, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rob62].

Jasani:1989:JO

[JR89] Bhupendra Jasani and Martin Rees. The junkyard in orbit. Bulletin of the Atomic Scientists, 45(8):24–25, 39, October 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [ER90].

Drollette:2016:FFS

[Jr16] Dan Drollette Jr. Fukushima future studies: five years later, what have we learned? Bulletin of the Atomic Scientists, 72(2): 97–104, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Drollette:2018:CHC

[Jr18] Dan Drollette Jr. A conversation with Helen Caldicott. Bulletin of the Atomic Scientists, 74(3):177–184, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1461976.

Drollette:2022:ISW

[Jr22a] Dan Drollette Jr. Interview with Sam West, founder of the Museum of Failure. Bulletin of the Atomic Scientists, 79(1):9–13, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2156685.

Drollette:2022:IRW

[Jr22b] Dan Drollette Jr. Introduction — Russia: what to expect next? Bulletin of the Atomic Scientists, 78(6):297–298, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2134527.

Drollette:2022:DKW

[Jr22c] Dan Drollette Jr. “It’s a different kind of world we’re living in now”: Interview with Francis Fukuyama. Bulletin of the Atomic Scientists, 78(6):318–323, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2132738.

Johansson:1979:SS

[JS79a] Thomas B. Johansson and Peter Steen. Solar Sweden. Bulletin of the Atomic Scientists, 35(8):19–22, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [GJS80].

Johansson:1979:WDR

[JS79b] Thomas B. Johansson and Peter Steen. What to do with radioactive waste? Bulletin of the Atomic Scientists, 35(9):38–42, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Jones:1997:AHH

[JS97] Rodney W. Jones and Nikolai N. Sokov. After Helsinki, the hard work. Bulletin of the Atomic Scientists, 53(4):26–30, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Janberg:2009:DCS

[JvH09] Klaus Janberg and Frank von Hippel. Dry-cask storage: How Germany led the way. Bulletin of the Atomic Scientists, 65 (5):24–32, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/65/5/24.

Jeppson:1963:LEO

[JvL63] Janet Jeppson and Theodore H. von Laue. Letter to the Editor: The Old Adam. Bulletin of the Atomic Scientists, 19(4):33–34, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vL63].

Kissinger:1988:LBU

[KA88] C. Clark Kissinger and Len Ackland. Letter: Bulletin’s untimely. Bulletin of the Atomic Scientists, 44(3):57–58, April 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack88h].

Kleinman:1990:LNP

[KA90] H. Kleinman and Len Ackland. Letter: Nuclear power series exercise in absurdity. Bulletin of the Atomic Scientists, 46(6):45, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack90j].

Kreitler:1991:LTM

[KA91] Walter M. Kreitler and Eric H. Arnett. Letters: Tomahawk missile fans on warpath. Bulletin of the Atomic Scientists, 47(6): 44–45, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Arn91].

Kaempffert:1953:NP

[Kae53] Waldeman Kaempffert. “Neither is perfect”. Bulletin of the Atomic Scientists, 9(9):332–333, 350, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaffka:1998:LJF

[Kaf98] Alexander Kaffka. Letter: Just the facts. Bulletin of the Atomic Scientists, 54(1):67, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Che97].

Kahler:1950:FPT

[Kah50] Erich Kahler. Foreign policy today. Bulletin of the Atomic Scientists, 6(12):356, 359–362, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kahn:1959:HMC

[Kah59] Herman Kahn. How many can be saved? Bulletin of the Atomic Scientists, 15(1):30–34, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kahn:1986:UMW

[Kah86] James S. Kahn. University management of weapons labs? Yes. Bulletin of the Atomic Scientists, 42(1):39–40, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See opposing view [Arc86].

Kahn:2007:ZC

[Kah07] Laura H. Kahn. The Zoonotic connection. Bulletin of the Atomic Scientists, 63(2):68, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/2/68.

Kalvin:1953:IFA

[Kal53] Harry Kalvin, Jr. Invoking the Fifth Amendment — some legal and practical considerations. Bulletin of the Atomic Scientists, 9 (5):181–184, 200, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1954:CJR

[Kal54] Harry Kalven, Jr. The case of J. Robert Oppenheimer before the Atomic Energy Commission. Bulletin of the Atomic Scientists, 10 (7):259–269, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Kal63b].

Kalven:1960:CTL

[Kal60] Harry Kalven, Jr. Congressional testing of Linus Pauling. Part I: The legal framework. Bulletin of the Atomic Scientists, 16(10): 383–389, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1961:CTL

[Kal61] Harry Kalven, Jr. Congressional testing of Linus Pauling. Part II: Sourwine in an old bottle. Bulletin of the Atomic Scientists, 17(1):12–19, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalkstein:1962:AT

[Kal62a] Marvin I. Kalkstein. On atmospheric testing. Bulletin of the Atomic Scientists, 18(3):34–35, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1962:BBP

[Kal62b] Harry Kalven, Jr. Books: The Price of Liberty, by Alan Barth. Bulletin of the Atomic Scientists, 18(5):39, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalkstein:1963:BBO

[Kal63a] Marvin Kalkstein. Books: Kill and Overkill: The Strategy of Annihilation, by Ralph E. Lapp. Bulletin of the Atomic Scientists, 19(6):26–27, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1963:CJR

[Kal63b] Harry Kalven, Jr. The case of J. Robert Oppenheimer before the Atomic Energy Commission. In Grodzins and Rabinowitch [GR63], pages 442–465. LCCN D842 .B78. Reprint of [Kal54].

Kalkstein:1964:PSN

[Kal64] Marvin Kalkstein. Preventing the spread of nuclear weapons. Bulletin of the Atomic Scientists, 20(10):18–19, December 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalkstein:1968:APD

[Kal68] Marvin Kalkstein. ABM: In the public domain: Book review: Debate the Antiballistic Missile, by Eugene Rabinowitch and Ruth Adams. Science, 160(3832):1098–1099, June 7, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/1723468.

Kallgren:1970:TPC

[Kal70] Joyce K. Kallgren. The third party in Chinese–American relations: The need for change. Bulletin of the Atomic Scientists, 26 (5):11–16, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaldor:1977:MTS

[Kal77a] Mary H. Kaldor. Military technology and social structure. Bulletin of the Atomic Scientists, 33(6):49–53, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kale:1977:LPW

[Kal77b] Roy A. Kale. Letter: People want peace. Bulletin of the Atomic Scientists, 33(10):3–4, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaldor:1978:BBA

[Kal78] Mary Kaldor. Books: The Arms Bazaar from Lebanon to Lockheed, by Anthony Sampson. Bulletin of the Atomic Scientists, 34 (4):58–59, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaldor:1981:CUK

[Kal81] Mary Kaldor. Commentary from the United Kingdom: END can be a beginning. Bulletin of the Atomic Scientists, 37(10):42– 46, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:BGZ

[Kal82a] Jamie Kalven. Bulletin: Ground Zero Week 1982. Bulletin of the Atomic Scientists, 38(4):57, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:BNW

[Kal82b] Jamie Kalven. Bulletin: Nuclear Weapons Freeze Campaign. Bulletin of the Atomic Scientists, 38(5):65, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:BFN

[Kal82c] Jamie Kalven. Bulletins: FAS Nuclear War/Education; FAS Petition Campaign; UCS/UCAM. Bulletin of the Atomic Scientists, 38(6):73, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:ESS

[Kal82d] Jamie Kalven. Election 1982: Security and secrecy. Bulletin of the Atomic Scientists, 38(8):16–17, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:NC

[Kal82e] Jamie Kalven. November 11 Convocation. Bulletin of the Atomic Scientists, 38(8):64, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:NW

[Kal82f] Jamie Kalven. Nuclear watch. Bulletin of the Atomic Scientists, 38(7):65, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:SP

[Kal82g] Jamie Kalven. Student Pugwash. Bulletin of the Atomic Scientists, 38(9):63–65, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1982:TLH

[Kal82h] Jamie Kalven. A talk with Louis Harris. Bulletin of the Atomic Scientists, 38(7):3–5, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sut83].

Kalven:1982:YRP

[Kal82i] Jamie Kalven. “Yellow Rain”: the public evidence. Bulletin of the Atomic Scientists, 38(5):15–20, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalter:1983:FET

[Kal83a] Joanmarie Kalter. The Fourth Estate in the Third World. Bulletin of the Atomic Scientists, 39(9):12–15, November 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1983:AVL

[Kal83b] Jamie Kalven. Atomic veterans: The legal quandary. Bulletin of the Atomic Scientists, 39(1):26–28, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [McK83].

Kalven:1983:GZA

[Kal83c] Jamie Kalven. Ground Zero announces war/peace game. Bulletin of the Atomic Scientists, 39(1):55, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaldor:1985:DES

[Kal85a] Mary Kaldor. Disengaging Europe from the superpowers. Bulletin of the Atomic Scientists, 41(7):140–143, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalet:1985:LMS

[Kal85b] Ira J. Kalet. Letter: More on strategic computing. Bulletin of the Atomic Scientists, 41(5):53, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Der85].

Kallenbach:1985:LST

[Kal85c] Ernst Kallenbach. Letter: Soviet testing flurry. Bulletin of the Atomic Scientists, 41(11):64, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dun85].

Kalven:1987:UVP

[Kal87] Jamie Kalven. U.S. visa policy: the machinery of exclusion. Bulletin of the Atomic Scientists, 43(4):21–30, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Per87a].

Kaldor:1988:ASE

[Kal88] Mary Kaldor. After the summit: End the Cold War in Europe. Bulletin of the Atomic Scientists, 44(2):26–28, March 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalven:1989:GSW

[Kal89] Jamie Kalven. Government secrecy: At war with the First Amendment. Bulletin of the Atomic Scientists, 45(1):56–59, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Swe89b].

Kaldor:1995:WKC

[Kal95a] Mary Kaldor. Who killed the Cold War. Bulletin of the Atomic Scientists, 51(4):57–60, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kalinowski:1995:LBT

[Kal95b] Martin Kalinowski. Letter: Ban tritium too. Bulletin of the Atomic Scientists, 51(1):65, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kamen:1946:AIB

[Kam46] Martin D. Kamen. The application of isotopes to biology. Bulletin of the Atomic Scientists, 1(12):9–11, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kamins:1971:BBC

[Kam71] Robert M. Kamins. Books: The Challenge of World Poverty: A World Anti-Poverty Program in Outline, by Gunnar Myrdal. Bulletin of the Atomic Scientists, 27(3):45–47, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kamen:1974:AOT

[Kam74] Martin D. Kamen. ‘All in our time’: The birthplace of Big Science. Bulletin of the Atomic Scientists, 30(9):42–46, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kamenetzky:1979:DP

[Kam79] Mario Kamenetzky. Development for the people. Bulletin of the Atomic Scientists, 35(6):44–46, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kamp:1996:ON

[Kam96] Karl-Heinz Kamp. An overrated nightmare. Bulletin of the Atomic Scientists, 52(4):30–34, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kamp:2005:PFF

[Kam05] Karl-Heinz Kamp. Preemption: Far from forsaken. Bulletin of the Atomic Scientists, 61(2):??, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/2/26.

Kammen:2014:SEI

[Kam14] Daniel M. Kammen. Solar energy innovation and Silicon Valley. Bulletin of the Atomic Scientists, 70(5):45–53, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kammen:2020:HWR

[Kam20] Daniel M. Kammen. Over the hump: Have we reached the peak of carbon emissions? Bulletin of the Atomic Scientists, 76 (5):256–262, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1806585.

Kammen:2023:WWH

[Kam23] Daniel M. Kammen. Why what happened to Oppenheimer then is relevant now. Bulletin of the Atomic Scientists, 79(4):267– 271, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2223087.

Kantrowitz:1969:TMC

[Kan69] Arthur Kantrowitz. The test: Meeting the challenge of new technology. Bulletin of the Atomic Scientists, 25(9):20–22, 48, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kantrowitz:1971:CRS

[Kan71a] Arthur Kantrowitz. Comment: The relevance of space. Bulletin of the Atomic Scientists, 27(4):32–33, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [KP71] and reply [Kan71b].

Kantrowitz:1971:CKR

[Kan71b] Arthur Kantrowitz. Communications: Kantrowitz replies. Bulletin of the Atomic Scientists, 27(10):3, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kan71a, KP71].

Kantrowitz:1977:SCE

[Kan77a] Arthur Kantrowitz. The Science Court Experiment: criticisms and responses. Bulletin of the Atomic Scientists, 33(4):44–48, 50, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pas77].

Kantrowitz:1977:SSC

[Kan77b] Arthur Kantrowitz. [Sidebar:] Specifications for the court. Bulletin of the Atomic Scientists, 33(4):45, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kannyo:1981:HRA

[Kan81] Edward Kannyo. Human rights in Africa. Bulletin of the Atomic Scientists, 37(10):14–19, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaneko:1996:JNN

[Kan96] Kumao Kaneko. Japan needs no umbrella. Bulletin of the Atomic Scientists, 52(2):46–51, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kang:2009:RNK

[Kan09] Jungmin Kang. Redirecting North Korea’s nuclear workers. Bulletin of the Atomic Scientists, 65(1):48–55, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/1/48.

Kang:2011:SKF

[Kan11] Jungmin Kang. South Korea in focus: The politics of spent fuel storage and disposal. Bulletin of the Atomic Scientists, 67 (3):51–58, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/3/51.

Kaplan:1960:CDE

[Kap60] Martin M. Kaplan. Communicable diseases and epidemics. Bulletin of the Atomic Scientists, 16(6):237–240, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1961:BBP

[Kap61] Martin M. Kaplan. Books: Einstein on Peace, by Otto Nathan and Heinz Norden, with a preface by Bertrand Russell. Bulletin of the Atomic Scientists, 17(7):291–292, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapitza:1962:FS

[Kap62] Peter L. Kapitza. The future of science. Bulletin of the Atomic Scientists, 18(4):3–7, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tuc62].

Kaplan:1966:SEA

[Kap66] Martin Kaplan. Social effects of animal diseases in developing countries. Bulletin of the Atomic Scientists, 22(9):15–21, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1967:BBP

[Kap67] Morton A. Kaplan. Books: Men of the Pentagon: From Forrestal to McNamara, by C. W. Borklund. Bulletin of the Atomic Scientists, 23(10):32–33, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapur:1972:IAB

[Kap72] Ashok Kapur. India after Bangladesh. Bulletin of the Atomic Scientists, 28(9):28–30, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapur:1974:IAB

[Kap74] Ashok Kapur. India and the atom: The blast was a signal to the Nuclear Club. Bulletin of the Atomic Scientists, 30(7):27– 30, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1976:HCB

[Kap76] Martin M. Kaplan. Health care and biomedical research in poor countries. Bulletin of the Atomic Scientists, 32(3):30–31, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1978:R

[Kap78a] Fred M. Kaplan. Response. Bulletin of the Atomic Scientists, 34 (4):51, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gou78c].

Kaplan:1978:SCDa

[Kap78b] Fred M. Kaplan. The Soviet civil defense myth. Bulletin of the Atomic Scientists, 34(3):14–20, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Whi78a, Haa78].

Kaplan:1978:SCDb

[Kap78c] Fred M. Kaplan. The Soviet civil defense myth: Part 2. Bulletin of the Atomic Scientists, 34(4):41–48, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplin:1979:CVP

[Kap79] Martin M. Kaplin. Commentary from Vienna: A personal reflection. Bulletin of the Atomic Scientists, 35(10):53–55, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1980:WCP

[Kap80] Martin M. Kaplan. What can Pugwash do? Bulletin of the Atomic Scientists, 36(9):13–14, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapitza:1981:GPI

[Kap81a] Peter Kapitza. Global problems, international solutions. Bulletin of the Atomic Scientists, 37(1):40–43, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Flo81].

Kaplan:1981:CNB

[Kap81b] Fred Kaplan. Commentary: The neutron bomb: What it is, the way it works. Bulletin of the Atomic Scientists, 37(8):6–7, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1982:YNW

[Kap82] Fred Kaplan. 37 years of nuclear weapons: Strategic thinkers. Bulletin of the Atomic Scientists, 38(10):51–56, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1983:LMD

[Kap83a] Daniel Kaplan. Lasers for missile defense. Bulletin of the Atomic Scientists, 39(5):5–8, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1983:AV

[Kap83b] Martin M. Kaplan. Another view. Bulletin of the Atomic Scientists, 39(9):27, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WS83].

Kaplan:1983:CP

[Kap83c] Martin M. Kaplan. Commentary: From Pugwash. Bulletin of the Atomic Scientists, 39(2):52, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapitza:1985:SVN

[Kap85] Sergei Kapitza. A Soviet view of nuclear winter. Bulletin of the Atomic Scientists, 41(9):37–39, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1989:LPP

[Kap89] Martin M. Kaplan. Letter: Pugwash promoted nonoffensive defense. Bulletin of the Atomic Scientists, 45(8):43, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapitsa:1990:SKR

[Kap90a] Sergei Kapitsa. Sergei Kapitsa remembers. Bulletin of the Atomic Scientists, 46(3):31, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1990:LCA

[Kap90b] Martin M. Kaplan. Letter: Chemical agreement highly disappointing. Bulletin of the Atomic Scientists, 46(7):45, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa90a] and comment [Bur90b].

Kapur:1990:DT

[Kap90c] Ashok Kapur. Dump the treaty. Bulletin of the Atomic Scientists, 46(6):21–22, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaplan:1991:BRP

[Kap91a] Cynthia S. Kaplan. Book review: the perestroika generation The Second Socialist Revolution, by Tatyana Zaslavskaya. Bulletin of the Atomic Scientists, 47(3):42, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kri91].

Kapur:1991:NWO

[Kap91b] Ashok Kapur. New world order: But will it play in Panmunjom? Bulletin of the Atomic Scientists, 47(5):36–37, June 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapitza:1992:PSS

[Kap92] Sergei Kapitza. Perspective: Soviet scientists: low pay, no pay, now insults. Bulletin of the Atomic Scientists, 48(4):8–9, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dow92].

Kaplan:1993:LMB

[Kap93] Martin M. Kaplan. Letter: McNamara, on balance. Bulletin of the Atomic Scientists, 49(7):55, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mea93].

Kapitza:1994:RSS

[Kap94] Sergei P. Kapitza. Russian science: Snubbed and sickly. Bulletin of the Atomic Scientists, 50(3):46–52, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kapur:1995:NWB

[Kap95] Ashok Kapur. Nonproliferation: Western biases. Bulletin of the Atomic Scientists, 51(1):38–43, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karber:1972:NPC

[Kar72] Phillip A. Karber. The Nixon policy on CBW. Bulletin of the Atomic Scientists, 28(1):22–25, 27, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karas:1980:BRB

[Kar80] Thomas H. Karas. Book review: The War Game: A Critique of Military Problem Solving, by Garry D. Brewer and Marvin Shubik. Bulletin of the Atomic Scientists, 36(2):49, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [HP80].

Karkoszka:1988:MJP

[Kar88a] Andrzej Karkoszka. Merits of the Jaruzelski plan. Bulletin of the Atomic Scientists, 44(7):32–34, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karp:1988:CWC

[Kar88b] Aaron Karp. Choosing what to control. Bulletin of the Atomic Scientists, 44(5):16, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karp:1988:FTW

[Kar88c] Aaron Karp. The frantic Third-World quest for ballistic missiles. Bulletin of the Atomic Scientists, 44(5):14–20, June 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karp:1988:TWB

[Kar88d] Aaron Karp. Third World ballistic missiles. Bulletin of the Atomic Scientists, 44(5):19, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [RC88].

Karl:1997:LDB

[Kar97] David J. Karl. Letter: Don’t blame India. Bulletin of the Atomic Scientists, 53(4):59–60, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BV97].

Karp:1999:UP

[Kar99] Aaron Karp. Uncertain prognosis. Bulletin of the Atomic Scientists, 55(1):66–69, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/66.

Karnad:2014:BBI

[Kar14] Bharat Karnad. Ban the bomb?: an Indian response. Bulletin of the Atomic Scientists, 70(6):14–17, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karako:2017:HMD

[Kar17] Thomas Karako. Homeland missile defense: How the United States got here. Bulletin of the Atomic Scientists, 73(3):159– 166, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1315035.

Kastler:1952:SFE

[Kas52] Alfred Kastler. Some French experiences. Bulletin of the Atomic Scientists, 8(7):242–243, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Kastenmeier:1960:LEB

[Kas60] Robert W. Kastenmeier. Letter to the Editor: Bouquets. Bulletin of the Atomic Scientists, 16(7):305, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kasdan:1964:TNO

[Kas64] Alan Richard Kasdan. Toward a new order in international trade. Bulletin of the Atomic Scientists, 20(10):40–43, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kash:1965:GSG

[Kas65] Don E. Kash. Is good scientists good politics? Bulletin of the Atomic Scientists, 21(3):34–36, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RMK65].

Kash:1967:TR

[Kas67] Don E. Kash. The tyranny of realism. Bulletin of the Atomic Scientists, 23(2):17–20, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kash:1969:FAS

[Kas69] Don E. Kash. Forces affecting science policy R&D. Bulletin of the Atomic Scientists, 25(4):10–15, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kastler:1975:BMP

[Kas75] Alfred Kastler. Bravo! Messieurs les Présidents: A Nobel laureate’s appeal to Ford and Brezhnev. Bulletin of the Atomic Scientists, 31(10):10–11, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kastler:1977:CCS

[Kas77] Alfred Kastler. The challenge of the century: Survival can only come by emptying the arsenals. Bulletin of the Atomic Scientists, 33(7):20–22, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Cultures 3(4) 1976. See comment [Hue77].

Kassof:1989:RC

[Kas89] Allen H. Kassof. The roots of change. Bulletin of the Atomic Scientists, 45(7):31, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khariton:1996:WIW

[KAS96] Yuli Khariton, Viktor Adamskii, and Yuri Smirnov. The way it was. Bulletin of the Atomic Scientists, 52(9):54–59, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/

1G1-18855733.html. See comment [HM97].

Kasperson:2012:SAR

[Kas12] Roger E. Kasperson. The social amplification of risk and low-level radiation. Bulletin of the Atomic Scientists, 68(3):59–66, May/ June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/68/3/59.

Kasperski:2015:NPU

[Kas15] Tatiana Kasperski. Nuclear power in Ukraine: Crisis or path to energy independence? Bulletin of the Atomic Scientists, 71(4): 43–50, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kassenova:2021:WBS

[Kas21] Togzhan Kassenova. Why Biden should push for ratification of the Comprehensive Test Ban Treaty. Bulletin of the Atomic Scientists, 77(1):13–14, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1859860.

Kimpton:1952:PAV

[KASN52] Lawrence A. Kimpton, Samuel K. Allison, Cyril S. Smith, and John U. Nef. A protest against visa situation. Bulletin of the Atomic Scientists, 8(1):4–5, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kate:1950:UEU

[Kat50a] J. J. Kate. Uranium exploration in the U.S. Bulletin of the Atomic Scientists, 6(6):177, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1950:UPP

[Kat50b] J. J. Katz. Uranium procurement policies. Bulletin of the Atomic Scientists, 6(11):349, 352, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1951:UR

[Kat51a] J. J. Katz. Uranium resources. Bulletin of the Atomic Scientists, 7(6):184, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1951:M

[Kat51b] Leonard I. Katzin. Manpower. Bulletin of the Atomic Scientists, 7(7–8):236–238, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1952:BRBa

[Kat52a] Leonard I. Katzin. Book review: Methods of Operations Research, by Philip M. Morse and George E. Kimball. Bulletin of the Atomic Scientists, 8(5):163, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1952:BRBb

[Kat52b] Leonard I. Katzin. Book review: Nobel — The Man and His Prizes, edited by the Nobel Foundation, and written by H.

Schück, R. Sohlman, A. Osterling, G. Liljestrand, A. Westgren, M. Siegbahn, A. Schou, and N. K. Stähl . Bulletin of the Atomic Scientists, 8(5):164, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1952:MUM

[Kat52c] Leonard I. Katzin. Manpower: Universal military training. Bulletin of the Atomic Scientists, 8(2):54–55, February 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1952:SM

[Kat52d] Leonard I. Katzin. Scientific manpower. Bulletin of the Atomic Scientists, 8(1):27–30, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1953:BBP

[Kat53a] Leonard I. Katzin. Books: Atomic Power: An Economic and Social Analysis, by Walter Isard and Vincent Whitney. Bulletin of the Atomic Scientists, 9(4):152, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1953:SPM

[Kat53b] Leonard I. Katzin. Scientific and professional manpower. Bulletin of the Atomic Scientists, 9(6):225–226, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1954:BBC

[Kat54a] L. I. Katzin. Books: The Common Sense of Science, by J. Bronowski. Bulletin of the Atomic Scientists, 10(6):232, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1954:LEL

[Kat54b] L. I. Katzin. Letter to the Editor: Lewis Mumford speech to the American Philosophical Society. Bulletin of the Atomic Scientists, 10(6):239, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1954:BBN

[Kat54c] Leonard Katzin. Books: Experimental Nuclear Physics, Volume I, edited by E. Segrè. Bulletin of the Atomic Scientists, 10(1): 30, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1954:BBB

[Kat54d] Leonard Katzin. Books: Isotopes in Biochemistry, edited by Davidson, Gray, and others. Bulletin of the Atomic Scientists, 10 (1):30, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1954:BBR

[Kat54e] Leonard Katzin. Books: Ultraviolet Radiation, by Lewis B. Koller. Bulletin of the Atomic Scientists, 10(6):238, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1954:BBF

[Kat54f] Leonard I. Katzin. Books: Energy in the Future, by Palmer Putnam. Bulletin of the Atomic Scientists, 10(3):98–99, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1957:BBE

[Kat57a] Leonard I. Katzin. Books: Atomic Energy and Congress, by Morgan Thomas. Bulletin of the Atomic Scientists, 13(4):145, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzin:1957:BBP

[Kat57b] Leonard I. Katzin. Books: Soviet Professional Manpower, by Nicholas DeWitt; Manpower and Education; The Growing Shortage of Scientists and Engineers; Shortage of Scientific and Engineering Manpower. Bulletin of the Atomic Scientists, 13(2):71– 72, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzenbach:1958:LLS

[Kat58] Nicholas Katzenbach. Law and lawyers in space. Bulletin of the Atomic Scientists, 14(6):220–224, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1961:ECH

[Kat61a] Amrom H. Katz. Events and comments: Hiders and finders. Bulletin of the Atomic Scientists, 17(10):423–424, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1961:STT

[Kat61b] Amrom H. Katz. Some things to think and some to do. Bulletin of the Atomic Scientists, 17(4):139–143, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klein:1971:FES

[KAT^{+}71] Marjorie H. Klein, A. A. Alexander, Kwo-Hwa Tseng, Milton H. Miller, Eng-Kung Yeh, and Hung-Ming Chu. Far Eastern students in a big university — subcultures within a subculture. Bulletin of the Atomic Scientists, 27(1):10, 16–19, January 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kathren:1973:CBR

[Kat73] Ronald L. Kathren. Communications: The BEIR report. Bulletin of the Atomic Scientists, 29(10):2, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tam73a, Tam73b].

Kathren:1975:LP

[Kat75] Ronald L. Kathren. Letters: Also present. Bulletin of the Atomic Scientists, 31(1):3, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author notes that Carl C. Gamertsfelder was omitted from the list of participants at the first test of the Chicago pile (CP-1) on 2 December 1942 given in [And74b, And74a].

Kates:1980:CMT

[Kat80a] Robert W. Kates. Commentary: Matters of truth. Bulletin of the Atomic Scientists, 36(5):11–12, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1980:PSI

[Kat80b] Arthur M. Katz. Perspective: SALT and international reconciliation. Bulletin of the Atomic Scientists, 36(5):3, May 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1990:CSP

[Kat90a] Mark N. Katz. Can the superpowers plot peace? Bulletin of the Atomic Scientists, 46(4):38–39, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Law90].

Katz:1990:SID

[Kat90b] Mark N. Katz. Soviet intervention didn’t pay. Bulletin of the Atomic Scientists, 46(4):39, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1992:RFP

[Kat92] Arthur Katz. Report: Fusion power: Is ITER the way? Bulletin of the Atomic Scientists, 48(8):8–10, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katz:1995:LCS

[Kat95] Jonathan Katz. Letter: Curatorship, not stewardship. Bulletin of the Atomic Scientists, 51(6):3, 72, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gus95, CB95].

Katsuta:2016:RTJ

[Kat16] Tadahiro Katsuta. The Rokkasho test: Has Japan learned the lessons of Fukushima? Bulletin of the Atomic Scientists, 72(3): 180–184, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kauffman:1980:CCN

[Kau80] Samuel Kauffman. Commentary: Clarification needed. Bulletin of the Atomic Scientists, 36(10):58, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kauffman:1982:CWS

[Kau82] Samuel Kauffman. Commentary: Wall of silence. Bulletin of the Atomic Scientists, 38(1):48, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaufmann:1990:PCM

[Kau90] William W. Kaufmann. A plan to cut military spending in half. Bulletin of the Atomic Scientists, 46(2):35–39, March 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kawakami:1978:NFC

[Kaw78] Kochi Kawakami. The nuclear fuel cycle in Japan. Bulletin of the Atomic Scientists, 34(6):17–18, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kaysen:1949:MIA

[Kay49] Carl Kaysen. Military importance of the atomic bomb. Bulletin of the Atomic Scientists, 5(12):340–343, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This is the US military’s response to comments in February 1949 issue on P. W. S. Blackett’s book Fear, War and the Bomb.

Kaysen:1966:AFS

[Kay66] Carl Kaysen. Allocating Federal support for basic research. Bulletin of the Atomic Scientists, 22(1):16–22, January 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kazmann:1961:LEM

[Kaz61] Raphael G. Kazmann. Letter to the Editor: Is McKinney consistent? Bulletin of the Atomic Scientists, 17(3):80, 121, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ham61c].

Kazi:2014:WWR

[Kaz14] Reshmi Kazi. Women and weapons: Redressing the gender gap: an Indian response. Bulletin of the Atomic Scientists, 70(5):8– 11, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khalilzad:1980:ENQ

[KB80] Zalmay Khalilzad and Cheryl Benard. Energy: no quick fix for a permanent crisis. Bulletin of the Atomic Scientists, 36(10):15– 20, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [vHWF^{+}81, BK81].

Kissinger:1988:FCC

[KB88] Henry Kissinger and Zbigniew Brzezinski. In favor of conventional counteroffensive: discriminate deterrence. Bulletin of the Atomic Scientists, 44(7):46, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kirby:1989:LAC

[KB89] Bernard C. Kirby and Louis A. P. Balázs. Letters: Arms control alternatives. Bulletin of the Atomic Scientists, 45(7):42–43, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelley:2003:MBB

[KC03] Marylia Kelley and Jay Coghlan. Mixing bugs and bombs. Bulletin of the Atomic Scientists, 59(5):24–31, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/5/24.

Kamen:1959:SNC

[KD59] Martin D. Kamen and Beka Doherty. Some new clichés about an old argument. Bulletin of the Atomic Scientists, 15(2):89– 91, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krass:1981:ESS

[KDFS81] Allan S. Krass, Christine T. Donovan, Lucia M. Ford, and Sandra C. Small. Energy self-sufficiency in a small city. Bulletin of the Atomic Scientists, 37(3):34–38, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ke:1972:CC

[Ke72] Tang Ke. The case of China. Bulletin of the Atomic Scientists, 28(7):54–55, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Tra72b].

Korb:2017:TOS

[KE17] Lawrence J. Korb and Carly Evans. The third offset strategy: A misleading slogan. Bulletin of the Atomic Scientists, 73(2):92–95, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1288443.

Kearny:1960:LES

[Kea60] Cresson H. Kearny. Letter to the Editor: Siting of ICBM bases. Bulletin of the Atomic Scientists, 16(10):374, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kearny:1961:RCC

[Kea61] Cresson H. Kearny. Reflections and comments: Controversy on Kahn. Bulletin of the Atomic Scientists, 17(7):293, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar61d, Sto61, Pet61, Mar61c].

Keane:2024:HDD

[Kea24] John Keane. How demagogues destroy democracy: a step-by-step global guide. Bulletin of the Atomic Scientists, 80(5):281– 286, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2388461.

Keeney:1956:LEH

[Kee56] Mrs. Philip Keeney. Letter to the Editor: On Harold Green’s review of William A. Reuben’s book The Atom Spy Hoax. Bulletin of the Atomic Scientists, 12(1):31–32, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keenleyside:1976:LNW

[Kee76] Hugh L. Keenleyside. Letter: Nuclear waste problem. Bulletin of the Atomic Scientists, 32(7):3–4, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Maw76] and comment [Maw77].

Keen:2007:SA

[Kee07] Sam Keen. A secular apocalypse. Bulletin of the Atomic Scientists, 63(1):29–31, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/29.

Kellogg:1961:RP

[Kel61] William W. Kellogg. Rocket probes. Bulletin of the Atomic Scientists, 17(5–6):199–203, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelly:1962:LED

[Kel62] George A. Kelly. Letter to the Editor: Deterrence as a stopgap. Bulletin of the Atomic Scientists, 18(3):38, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Pol62a].

Kellogg:1964:SUP

[Kel64] Charles E. Kellogg. Soil-use planning for individual and public goals. Bulletin of the Atomic Scientists, 20(9):15–18, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kellogg:1978:MWE

[Kel78] William W. Kellogg. Is mankind warming Earth. Bulletin of the Atomic Scientists, 34(2):10–19, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelly:1979:BRB

[Kel79] Henry Kelly. Book reviews: Principle Conclusions of the American Physical Society Study Group on Solar Photovoltaic Energy Conversion, by Henry Ehrenreich, Chairman. Bulletin of the Atomic Scientists, 35(4):36–38, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keller:1984:BND

[Kel84] Abraham C. Keller. Breaking the nuclear deadlock: Ten steps to peace. Bulletin of the Atomic Scientists, 40(3):51–52, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelly:1987:LOF

[Kel87] Grahame Kelly. Letter: Old-fashioned wisdom. Bulletin of the Atomic Scientists, 43(7):58, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi87c].

Kelly:1990:LHG

[Kel90] Grahame J. Kelly. Letter: How green will the greenhouse be? Bulletin of the Atomic Scientists, 46(9):46, November 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi90a].

Kellman:1998:BRH

[Kel98a] Barry Kellman. Book review: Hi-tech pestilence: The Eleventh Plague: The Politics of Biological and Chemical Warfare, by Leonard A. Cole. Bulletin of the Atomic Scientists, 54(4):62–64, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelly:1998:LSP

[Kel98b] Wes Kelly. Letter: Some plane truth. Bulletin of the Atomic Scientists, 54(4):4–5, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mat98e].

Kellett:2003:BF

[Kel03] Frederick Kellett. Bankrolling failure. Bulletin of the Atomic Scientists, 59(5):42–45, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/5/42.

Kelley:2010:BNP

[Kel10] Robert Kelley. Burma nuclear program highlights the need for a standing UN technical body. Bulletin of the Atomic Scientists, 66(6):67–76, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/66/6/67.

Kellogg:2020:MWE

[Kel20] William W. Kellogg. 1978: Is mankind warming the Earth? Bulletin of the Atomic Scientists, 76(6):381–390, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847497.

Kemper:1997:LFT

[Kem97] William A. Kemper. Letter: Forget two tracks: store it. Bulletin of the Atomic Scientists, 53(2):60–61, March/April 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kemburi:2009:TNR

[Kem09] Kalyan Kemburi. A Taiwanese nuclear revival. Bulletin of the Atomic Scientists, 65(2):13–22, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/2/13.

Kennan:1953:CC

[Ken53] George F. Kennan. Communism and conformity. Bulletin of the Atomic Scientists, 9(8):296–298, 304, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1960:WGP

[Ken60a] John F. Kennedy. As we go to press: Kennedy opposes resumption of atom tests now. Bulletin of the Atomic Scientists, 16(9): 377, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter to Thomas E. Murray, former member of the Atomic Energy Commission.

Kennedy:1960:DCW

[Ken60b] John F. Kennedy. Disarmament Can be won. Bulletin of the Atomic Scientists, 16(6):217–219, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1961:KNF

[Ken61a] John F. Kennedy. The Kennedy new frontier. Bulletin of the Atomic Scientists, 17(3):91, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1961:MRS

[Ken61b] John F. Kennedy. Messages received at Stowe: From the President of the U.S. Bulletin of the Atomic Scientists, 17(9):385, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1961:RCW

[Ken61c] John F. Kennedy. Reflections and comments: “We need your help” — President Kennedy. Bulletin of the Atomic Scientists, 17(8):344, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1962:MBA

[Ken62] John F. Kennedy. Message to the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 18(10):2, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1963:IC

[Ken63a] John F. Kennedy. International cooperation. Bulletin of the Atomic Scientists, 19(10):42, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1963:KU

[Ken63b] John F. Kennedy. Kennedy at the U.N. Bulletin of the Atomic Scientists, 19(9):45, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennett:1963:PNN

[Ken63c] Lord Kennett. [Proliferation of national nuclear forces]. Bulletin of the Atomic Scientists, 19(4):11, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1965:A

[Ken65] John F. Kennedy. [The atom]. Bulletin of the Atomic Scientists, 21(1):49, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1966:APS

[Ken66] Robert F. Kennedy. The Alliance for Progress: Symbol and substance. Bulletin of the Atomic Scientists, 22(9):28–34, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1971:SSE

[Ken71] Edward M. Kennedy. Statement of Senator Edward M. Kennedy on re-employing defense scientists and engineers. Bulletin of the Atomic Scientists, 27(3):41–43, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kent:1977:LGR

[Ken77] Rudolph J. Kent. Letter: The golden rule. Bulletin of the Atomic Scientists, 33(5):2–3, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Agn77a].

Kennan:1978:ARS

[Ken78] George F. Kennan. America’s relations with the Soviet Union. Bulletin of the Atomic Scientists, 34(6):21–23, June 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kendall:1979:PIF

[Ken79] Henry W. Kendall. Panel III of the First Nuclear War Conference: Second strike. Bulletin of the Atomic Scientists, 35(7):32–37, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennan:1980:CIR

[Ken80] George F. Kennan. Commentary: Imprudent response to the Afghan crisis? Bulletin of the Atomic Scientists, 36(4):7–9, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennan:1981:SPE

[Ken81] George F. Kennan. Scholarship, politics and the East–West relationship. Bulletin of the Atomic Scientists, 37(5):4–7, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kenworthy:1983:WUS

[Ken83] Eldon Kenworthy. Why the United States is in Central America. Bulletin of the Atomic Scientists, 39(8):14–18, 20, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SL83, Wio84, Mos84].

Kennet:1985:SWE

[Ken85] Wayland Kennet. Star Wars: Europe’s polite waffle. Bulletin of the Atomic Scientists, 41(8):7–11, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SK85].

Kennan:1987:KR

[Ken87a] George Kennan. Kennan redux, 1987. Bulletin of the Atomic Scientists, 43(9):22, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennan:1987:XC

[Ken87b] George F. Kennan. “X” on containment, 1947. Bulletin of the Atomic Scientists, 43(9):16, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennet:1987:LBP

[Ken87c] Wayland Kennet. Letter: British party politics and nuclear weapons. Bulletin of the Atomic Scientists, 43(2):60–61, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gap86].

Kennedy:2007:ZT

[Ken07] John F. Kennedy. Zero tolerance. Bulletin of the Atomic Scientists, 63(1):84, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/84.

Kepner:1976:LSE

[Kep76] Theodore S. Kepner. Letter: Safety over-emphasis. Bulletin of the Atomic Scientists, 32(2):2–3, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keppel:1982:CCF

[Kep82] David Keppel. Commentary: Cheating on the freeze — not the issue. Bulletin of the Atomic Scientists, 38(9):60, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kerrigan:1954:BBJ

[Ker54] Anthony Kerrigan. Books: Understanding the Japanese Mind, by James C. Moloney. Bulletin of the Atomic Scientists, 10(6):232, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kerr:1968:BBS

[Ker68] Malcolm H. Kerr. Books: Education and Science in the Arab World, by Fahim I. Qubain. Bulletin of the Atomic Scientists, 24 (3):30, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kerker:1976:LLS

[Ker76] Milton Kerker. Letter: Lack of sensibility. Bulletin of the Atomic Scientists, 32(9):5, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Kerr:2006:DW

[Ker06] Paul Kerr. Divided from within. Bulletin of the Atomic Scientists, 62(6):17–19, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/6/17.

Kessler:1985:PUP

[Kes85a] Richard J. Kessler. The Philippines: a U.S. policy dilemma. Bulletin of the Atomic Scientists, 41(1):41–44, January 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kessler:1985:R

[Kes85b] Richard J. Kessler. A rejoinder. Bulletin of the Atomic Scientists, 41(5):40, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha85a].

Kessler:1989:PSN

[Kes89] Richard Kessler. Peronists seek “nuclear greatness”. Bulletin of the Atomic Scientists, 45(4):13–15, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keulman:1983:BRB

[Keu83] Kenneth Keulman. Book review: Indefensible Weapons: The Political and Psychological Case Against Nuclearism, by Robert Jay Lifton and Richard Falk. Bulletin of the Atomic Scientists, 39(4):35, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kevles:2007:WSG

[Kev07] Daniel J. Kevles. When science goes bad. Bulletin of the Atomic Scientists, 63(1):73–75, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/1/73.

Keys:1957:LES

[Key57] Donald Fraser Keys. Letter to the Editor: Spirit and matter. Bulletin of the Atomic Scientists, 13(3):113, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keyfitz:1966:PPT

[Key66] Nathan Keyfitz. Privilege and poverty: Two worlds on one planet. Bulletin of the Atomic Scientists, 22(3):9–14, March 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keyfitz:1980:CCS

[Key80] Nathan Keyfitz. Commentary: Crossroads for science. Bulletin of the Atomic Scientists, 36(6):7–8, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Keynan:1981:BRB

[Key81] Alex Keynan. Book review: Advice to a Young Scientist, by Peter B. Medawar. Bulletin of the Atomic Scientists, 37(3):45, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1950:BRB

[KFM50] Peter Kihss, Robert W. Frase, and Donald Meiklejohn. Book review: Minutes to Midnight: The International Control of Atomic Energy, edited, with commentary, by Eugene Rabinowitch. Bulletin of the Atomic Scientists, 6(11):344–347, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Korneichuk:1961:SAS

[KFT61] A. Korneichuk, E. Fedorov, and N. Talensky. Soviets answer SANE. Bulletin of the Atomic Scientists, 17(7):295–296, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kopp:2025:CCW

[KGS25] Robert Kopp, Elisabeth Gilmore, and Rachael Shwom. Climate change will surprise us, but so-called ‘tipping points’ may lead us astray. Bulletin of the Atomic Scientists, 81(2):121–125, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2025.2464445.

Killian:1954:CD

[KH54] James R. Killian, Jr. and A. G. Hill. For a continental defense. Bulletin of the Atomic Scientists, 10(1):13–16, 30, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

King-Hall:1963:LEU

[KH63] Stephen King-Hall. Letter to the Editor: Unilateral disarmament. Bulletin of the Atomic Scientists, 19(5):34–35, May 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knoll:1985:LC

[KH85] Erwin Knoll and Morton Halperin. Letter: A clarification. Bulletin of the Atomic Scientists, 41(10):55, November 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hal85].

Kurzrok:2013:BCI

[KH13] Andrew Kurzrok and Gretchen Hund. Beyond compliance: Integrating nonproliferation into corporate sustainability. Bulletin of the Atomic Scientists, 69(3):31–42, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/3/31.

Khalilzad:1980:PB

[Kha80] Zalmay Khalilzad. Pakistan and the bomb. Bulletin of the Atomic Scientists, 36(1):11–16, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khan:1998:PJC

[Kha98] Ayesha Khan. Pakistan joins the club. Bulletin of the Atomic Scientists, 54(4):34–39, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khan:1999:NN

[Kha99a] Ayesha Khan. Nuclear but needy. Bulletin of the Atomic Scientists, 55(1):12, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khan:1999:PNN

[Kha99b] Ayesha Khan. Pakistan: Nuclear but needy. Bulletin of the Atomic Scientists, 55(1):12–13, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/12.

Kang:2005:SKN

[KHB^{+}05] Jungmin Kang, Peter Hayes, Li Bin, Tatsujiro Suzuki, and Richard Tanter. South Korea’s nuclear surprise. Bulletin of the Atomic Scientists, 61(1):40–49, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/40.

Khrushchev:1957:KSE

[Khr57] Nikita Khrushchev. Khrushchev speaks on economic and technical progress. Bulletin of the Atomic Scientists, 13(10):360, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khrushchev:1961:MRS

[Khr61] Nikita Khrushchev. Messages received at Stowe: From the Premier of the USSR. Bulletin of the Atomic Scientists, 17(9):385, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khrushchev:1963:KOE

[Khr63] Nikita Khrushchev. Khrushchev on overkill: East Berlin, January 16, 1963. Bulletin of the Atomic Scientists, 19(4):7, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khripunov:1994:RRD

[Khr94] Igor Khripunov. Report: Russia: Delusions v. conversion. Bulletin of the Atomic Scientists, 50(4):11–13, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khripunov:1996:PP

[Khr96a] Igor Khripunov. Perspective: The parity’s over. Bulletin of the Atomic Scientists, 52(5):16–19, September/October 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khripunov:1996:RRE

[Khr96b] Igor Khripunov. Report: Russian election: Red army blues. Bulletin of the Atomic Scientists, 52(3):13–14, 55, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khripunov:1997:GWT

[Khr97] Igor Khripunov. Have guns, will travel. Bulletin of the Atomic Scientists, 53(3):47–51, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ise97].

Khripunov:1999:ME

[Khr99a] Igor Khripunov. Minatom at the edge. Bulletin of the Atomic Scientists, 55(3):56–61, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/3/56.

Khripunov:1999:MR

[Khr99b] Igor Khripunov. Moscow reacts. Bulletin of the Atomic Scientists, 55(4):32–35, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/4/32.

Khripunov:2000:LLT

[Khr00] Igor Khripunov. Last leg of the triad. Bulletin of the Atomic Scientists, 56(4):58–64, July/August 2000. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/4/58.

Khripunov:2005:NSA

[Khr05] Igor Khripunov. Nuclear security: Attitude check. Bulletin of the Atomic Scientists, 61(1):58–64, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/58.

Khripunov:2010:IPN

[Khr10] Igor Khripunov. Involving the public in nuclear terrorism preparedness: Reflections on the Washington nuclear summit. Bulletin of the Atomic Scientists, 66(4):33–41, July/August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/4/33.

Kianitsa:1993:TA

[Kia93] Victor Kianitsa. Test anxiety. Bulletin of the Atomic Scientists, 49(8):37–39, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano93-30].

Kibaroglu:2001:GD

[Kib01] Mustafa Kibaroglu. The Generals’ discontent. Bulletin of the Atomic Scientists, 57(2):28–30, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/28.

Kibaroglu:2003:TSN

[Kib03] Mustafa Kibaroglu. Turkey says no. Bulletin of the Atomic Scientists, 59(4):22–25, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/4/22.

Kibaroglu:2004:TMB

[Kib04] Mustafa Kibaroğlu. Turkey: Missing Bill Clinton. Bulletin of the Atomic Scientists, 60(2):30–32, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/30.

Kibaroglu:2007:TNT

[Kib07] Mustafa Kibaroglu. A Turkish nuclear turnaround. Bulletin of the Atomic Scientists, 63(6):64, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/6/64.

Kibaroglu:2010:IQH

[Kib10] Mustafa Kibaroglu. The Iranian quagmire: How to move forward: Position: Resuscitate the nuclear swap deal. Bulletin of the Atomic Scientists, 66(6):102–108, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/6/102.

Kibaroglu:2012:HCC

[Kib12] Mustafa Kibaroglu. How can countries ensure that the nuclear security summit does not lose momentum and become just another gathering?: Position: Kickstart momentum with local review teams and summits with teeth. Bulletin of the Atomic Scientists, 68(2):81–83, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/2/81.

Kibaroglu:2017:BBB

[Kib17] Mustafa Kibaroglu. Ban the bomb by … banning the bomb? Bulletin of the Atomic Scientists, 73(3):199–200, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1315107.

Kidderr:1991:NWS

[Kid91] Ray E. Kidderr. Nuclear weapon safety: Safety no barrier to test ban. Bulletin of the Atomic Scientists, 47(3):32–34, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1947:AENa

[Kih47a] Peter Kihss. UN atomic energy news. Bulletin of the Atomic Scientists, 3(7):193–194, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1947:AENb

[Kih47b] Peter Kihss. UN atomic energy news. Bulletin of the Atomic Scientists, 3(9):247–251, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1947:UNAh

[Kih47c] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 3(8):213–215, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1947:UNAj

[Kih47d] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 3(10):295–298, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1947:UNAl

[Kih47e] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 3(12):374–376, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1947:UNAk

[Kih47f] Peter Kihss. United Nations atomic energy news: Mr. Aranha opens the Assembly; U.S. and Poland on disarmament; comment on progress of A.E.C.; blames U.S for failure; names warmongers and offers resolution; small countries express opinions; comments by Mr. Austen and Mr. McNeil; atom bomb an “Unfortunate discovery”; fear and suspicion real cause of delay. Bulletin of the Atomic Scientists, 3(11):329–331, 334, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAa

[Kih48a] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(1):18–20, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAb

[Kih48b] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(2):55–56, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAc

[Kih48c] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(3):83–85, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAd

[Kih48d] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(4):125–127, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAe

[Kih48e] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(5):159–160, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAf

[Kih48f] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(6):189–190, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAg

[Kih48g] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(7):219–221, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAh

[Kih48h] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(8):253–255, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAi

[Kih48i] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(9):286–287, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1948:UNAj

[Kih48j] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 4(10):318–320, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kihss:1949:UNA

[Kih49] Peter Kihss. United Nations atomic energy news. Bulletin of the Atomic Scientists, 5(1):29–31, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Killian:1957:STT

[Kil57] James R. Killian, Jr. [Scientists on tap but not on top?]. Bulletin of the Atomic Scientists, 13(2):70, February 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the American Scientist, April 1956.

Killian:1959:SPP

[Kil59] James R. Killian, Jr. Science and public policy. Bulletin of the Atomic Scientists, 15(4):168–172, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Killian:1962:SER

[Kil62] James R. Killian, Jr. Science and engineering: Resources for peace. Bulletin of the Atomic Scientists, 18(3):2–5, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Killian:1965:SSD

[Kil65] J. R. Killian, Jr. Science in the State Department: A practical imperative. Bulletin of the Atomic Scientists, 21(5):12–17, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Killian:1978:BRB

[Kil78] James R. Killian, Jr. Book reviews: The Apes to Warlords: the Autobiography (1904–1946) of Solly Zuckerman, by Hamish Hamilton. Bulletin of the Atomic Scientists, 34(10):41–44, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kimura:1987:CAR

[Kim87] Motoharu Kimura. Choice & actions: Reflections of a Japanese physicist. Bulletin of the Atomic Scientists, 43(9):7–10, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kim:1995:LSK

[Kim95] Taewoo Kim. Letter: South Korean patience wearing thin. Bulletin of the Atomic Scientists, 51(5):3, 59, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AO95c].

Kimball:2001:ETB

[Kim01] Daryl Kimball. End the test ban limbo. Bulletin of the Atomic Scientists, 57(1):30–31, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/30.

Kimmerling:2006:STI

[Kim06] Baruch Kimmerling. Seeing through it all. Bulletin of the Atomic Scientists, 62(2):59–60, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/2/59.

Kim:2020:HKS

[Kim20] Duyeon Kim. How to keep South Korea from going nuclear. Bulletin of the Atomic Scientists, 76(2):68–75, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1728966.

Kim:2022:EAI

[Kim22] Lami Kim. Exchanging atoms for influence: Competition in Southeast Asia’s nuclear market. Bulletin of the Atomic Scientists, 78(2):84–90, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2038890.

King:1954:TAU

[Kin54] J. R. King. Technical assistance in U.S. foreign policy. Bulletin of the Atomic Scientists, 10(4):127–128, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

King:1957:BMT

[Kin57] James E. King, Jr. Britain’s moment of truth: What it means to the U.S. Bulletin of the Atomic Scientists, 13(6):212–215, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kindleberger:1961:RD

[Kin61] Betty Kindleberger. Regional disarmament. Bulletin of the Atomic Scientists, 17(7):295, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

King:1967:LEC

[Kin67] Robert W. King. Letter to the Editor: Can our fractured world be mended? Bulletin of the Atomic Scientists, 23(1):31–32, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

King:1969:BBR

[Kin69] David Burnett King. Books: Weimar and the Rise of Hitler, by A. J. Nicholls. Bulletin of the Atomic Scientists, 25(10):41– 42, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

King:1974:LNR

[Kin74] Peter King. Letter: Non-retaliation. Bulletin of the Atomic Scientists, 30(7):3, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor74].

Kincade:1978:CWB

[Kin78a] William H. Kincade. Commentary: From Washington: Banning nuclear tests: cold feet in the Carter administration. Bulletin of the Atomic Scientists, 34(9):8–11, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1978:SAS

[Kin78b] William H. Kincade. A strategy for all seasons: targeting doctrine and strategic arms control. Bulletin of the Atomic Scientists, 34 (5):14–20, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1980:CAC

[Kin80a] William Kincade. Commentary: On the Arms Control Association. Bulletin of the Atomic Scientists, 36(3):63, March 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Div79].

Kincade:1980:BRBb

[Kin80b] William H. Kincade. Book review: Politics of Arms Control: The Role and Effectiveness of the U.S. Arms Control and Disarmament Agency, by Duncan L. Clarke. Bulletin of the Atomic Scientists, 36(8):50–54, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1980:BRBa

[Kin80c] William H. Kincade. Book review: The Contaminant, by Leonard Reiffel. Bulletin of the Atomic Scientists, 36(6):50, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1981:CWA

[Kin81a] William H. Kincade. Commentary from Washington: Another summer of the MX. Bulletin of the Atomic Scientists, 37(7):50– 51, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1981:CWR

[Kin81b] William H. Kincade. Commentary from Washington: Reagan and SALT: the rhetoric and the reality. Bulletin of the Atomic Scientists, 37(6):9–10, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1983:WCE

[Kin83] William H. Kincade. War concerns in Europe. Bulletin of the Atomic Scientists, 39(1):46–48, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1984:AAC

[Kin84a] William H. Kincade. Agenda 1984: Arms control priorities. Bulletin of the Atomic Scientists, 40(1):6–7, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

King:1984:CCW

[Kin84b] Ashley King. Commentary: Can war be avoided? Bulletin of the Atomic Scientists, 40(8):49, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincaide:1986:BRB

[Kin86] William H. Kincaide. Book reviews: Negotiating from Strength: Leverage in U.S.-Soviet Arms Control Negotiations , by Robert J. Einhorn; A Game for High Stakes: Lessons Learned in Negotiating with the Soviet Union, edited by Leon Sloss and M. Scott Davis. Bulletin of the Atomic Scientists, 42(6):54–56, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kirk:1955:LEH

[Kir55] John R. Kirk. Letter to the Editor: [H-bomb threats]. Bulletin of the Atomic Scientists, 11(10):380, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Say55, SD55].

Kiron:1963:NII

[Kir63] Allan Kiron. New incentives for the inventor. Bulletin of the Atomic Scientists, 19(10):13, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kireyev:1990:MAM

[Kir90] Alexei Kireyev. Ministry of Aviation and Macaroni? Bulletin of the Atomic Scientists, 46(1):18, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kirstein:1991:LWD

[Kir91] Peter N. Kirstein. Letter: Wrong dates. Bulletin of the Atomic Scientists, 47(6):45–46, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa91f].

Kirby:2017:TAM

[Kir17] Reid Kirby. The Trump’s administration’s misaligned approach to national biodefense. Bulletin of the Atomic Scientists, 73 (6):382–387, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1388674.

Kishimura:1954:LEB

[Kis54] Kinjiro Kishimura. Letter to the Editor: The Bikini bomb tests. Bulletin of the Atomic Scientists, 10(8):335, 337, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1960:SFA

[Kis60] George Kistiakowsky. Science and foreign affairs. Bulletin of the Atomic Scientists, 16(4):114–116, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1966:ASB

[Kis66] G. B. Kistiakowsky. Allocating support for basic research — and the importance of practical applications. Bulletin of the Atomic Scientists, 22(2):12–18, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowski:1969:ORS

[Kis69] George Kistiakowski. Other reactions in the science community. Bulletin of the Atomic Scientists, 25(6):48, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1972:ASC

[Kis72] G. B. Kistiakowsky. American science at the crossroads. Bulletin of the Atomic Scientists, 28(7):4–7, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1976:SDM

[Kis76a] G. B. Kistiakowsky. Are several decision-makers better than one? Bulletin of the Atomic Scientists, 32(3):55–57, March 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1976:NFD

[Kis76b] G. B. Kistiakowsky. A new form of diffuse responsibility? Bulletin of the Atomic Scientists, 32(3):58, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sto76a].

Kistiakowsky:1978:FNB

[Kis78] George Kistiakowsky. The folly of the neutron bomb. Bulletin of the Atomic Scientists, 34(7):25–29, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1979:CGB

[Kis79a] G. B. Kistiakowsky. Commentary: The good and the bad of nuclear arms control negotiations. Bulletin of the Atomic Scientists, 35(5):7–9, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1979:PHN

[Kis79b] George Kistiakowsky. Panel I: How a nuclear war can start … from an unrestrained arms race. Bulletin of the Atomic Scientists, 35(4):23–24, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1980:TR

[Kis80] George B. Kistiakowsky. Trinity — a reminiscence. Bulletin of the Atomic Scientists, 36(6):19–22, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1981:CNY

[Kis81] George B. Kistiakowsky. Commentary from New York: Carrying the message. Bulletin of the Atomic Scientists, 37(3):10, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1982:FA

[Kis82] George B. Kistiakowsky. The four anniversaries. Bulletin of the Atomic Scientists, 38(10):2–3, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kistiakowsky:1989:PKP

[Kis89] Vera Kistiakowsky. Perspective: Keep Pentagon out of civilian economy. Bulletin of the Atomic Scientists, 45(3):5, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kiselyov:1993:OMM

[Kis93a] Sergei Kiselyov. Our man in Missouri. Bulletin of the Atomic Scientists, 49(2):34–38, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kiselyov:1993:PNC

[Kis93b] Sergei Kiselyov. Perspective: Nothing in common, no wealth. Bulletin of the Atomic Scientists, 49(1):12–13, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kiselyov:1993:RUV

[Kis93c] Sergei Kiselyov. Report: Ukraine: The view from Kiev. Bulletin of the Atomic Scientists, 49(9):6–8, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kiselyov:1993:USG

[Kis93d] Sergei Kiselyov. Ukraine: Stuck with the goods. Bulletin of the Atomic Scientists, 49(2):30–33, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kiselyov:1994:UWA

[Kis94] Sergei Kiselyov. Ukraine: Not so Western after all. Bulletin of the Atomic Scientists, 50(1):32–35, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kiselyov:1996:IBC

[Kis96] Sergei Kiselyov. Inside the beast. Bulletin of the Atomic Scientists, 52(3):43–51, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translation from Russian by Viktoria Tripolskaya-Mitlyng.

Krane:2016:NEM

[KJE16] Jim Krane, Amy Myers Jaffe, and Jareer Elass. Nuclear energy in the Middle East: Chimera or solution? Bulletin of the Atomic Scientists, 72(1):44–51, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kennedy:1963:TBK

[KK63] John F. Kennedy and Nikita Khrushchev. Test ban: Kennedy– Khrushchev exchange. Bulletin of the Atomic Scientists, 19(3): 32–37, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Katzir-Katchalsky:1972:ISA

[KK72] Aharon Katzir-Katchalsky. An Israeli scientist’s approach to human values. Bulletin of the Atomic Scientists, 28(8):19–23, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Koblitz:1985:LSC

[KK85] Neal Koblitz and Kevin Klose. Letter: Soviet censorship. Bulletin of the Atomic Scientists, 41(10):57, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Klo85].

Krepinevich:1998:APS

[KK98a] Andrew F. Krepinevich and Steven M. Kosiak. Advances in precision-strike capabilities. Bulletin of the Atomic Scientists, 54(6):31, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepinevich:1998:SBF

[KK98b] Andrew F. Krepinevich and Steven M. Kosiak. Smarter bombs, fewer nukes. Bulletin of the Atomic Scientists, 54(6):26–32, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khripunov:2002:RNI

[KK02] Igor Khripunov and Maria Katsva. Russia’s nuclear industry: The next generation. Bulletin of the Atomic Scientists, 58(2): 51–57, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/2/51.

Kim:2012:WNS

[KK12] Duyeon Kim and Jungmin Kang. Where nuclear safety and security meet. Bulletin of the Atomic Scientists, 68(1):86–93, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/1/86.

Kittner:2018:BIC

[KK18a] Noah Kittner and Daniel M. Kammen. A battery of innovative choices — if we commit to investing. Bulletin of the Atomic Scientists, 74(1):7–10, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1413224.

Kristensen:2018:INF

[KK18b] Hans M. Kristensen and Matt Korda. Indian nuclear forces, 2018. Bulletin of the Atomic Scientists, 74(6):361–366, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1533162.

Kristensen:2019:UBM

[KK19a] Matt Korda and Hans M. Kristensen. US ballistic missile defenses, 2019. Bulletin of the Atomic Scientists, 75(6):295–306, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1680055.

Korda:2019:CNF

[KK19b] Hans M. Kristensen and Matt Korda. Chinese nuclear forces, 2019. Bulletin of the Atomic Scientists, 75(4):171–178, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1628511.

Kristensen:2019:FNF

[KK19c] Hans M. Kristensen and Matt Korda. French nuclear forces, 2019. Bulletin of the Atomic Scientists, 75(1):51–55, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1556003.

Korda:2019:RNF

[KK19d] Hans M. Kristensen and Matt Korda. Russian nuclear forces, 2019. Bulletin of the Atomic Scientists, 75(2):73–84, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1580891.

Korda:2019:TNW

[KK19e] Hans M. Kristensen and Matt Korda. Tactical nuclear weapons, 2019. Bulletin of the Atomic Scientists, 75(5):252–261, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1654273.

Korda:2019:USN

[KK19f] Hans M. Kristensen and Matt Korda. United States nuclear forces, 2019. Bulletin of the Atomic Scientists, 75(3):122–134, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1606503.

Kristensen:2020:CNF

[KK20a] Hans M. Kristensen and Matt Korda. Chinese nuclear forces, 2020. Bulletin of the Atomic Scientists, 76(6):443–457, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1846432.

Kristensen:2020:INF

[KK20b] Hans M. Kristensen and Matt Korda. Indian nuclear forces, 2020. Bulletin of the Atomic Scientists, 76(4):217–225, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1778378.

Kristensen:2020:RNF

[KK20c] Hans M. Kristensen and Matt Korda. Russian nuclear forces, 2020. Bulletin of the Atomic Scientists, 76(2):102–117, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1728985.

Korda:2020:USN

[KK20d] Hans M. Kristensen and Matt Korda. United States nuclear forces, 2020. Bulletin of the Atomic Scientists, 76(1):46–60, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1701286.

Kristensen:2021:CNW

[KK21a] Hans M. Kristensen and Matt Korda. Chinese nuclear weapons, 2021. Bulletin of the Atomic Scientists, 77(6):318–336, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1989208.

Kristensen:2021:NKN

[KK21b] Hans M. Kristensen and Matt Korda. North Korean nuclear weapons, 2021. Bulletin of the Atomic Scientists, 77(4):222–236, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1940803.

Kristensen:2021:PNW

[KK21c] Hans M. Kristensen and Matt Korda. Pakistani nuclear weapons, 2021. Bulletin of the Atomic Scientists, 77(5):265–278, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1964258.

Kristensen:2021:RNW

[KK21d] Hans M. Kristensen and Matt Korda. Russian nuclear weapons, 2021. Bulletin of the Atomic Scientists, 77(2):90–108, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1885869.

Kristensen:2021:UKN

[KK21e] Hans M. Kristensen and Matt Korda. United Kingdom nuclear weapons, 2021. Bulletin of the Atomic Scientists, 77(3): 153–158, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1912309.

Kristensen:2021:USN

[KK21f] Hans M. Kristensen and Matt Korda. United States nuclear weapons, 2021. Bulletin of the Atomic Scientists, 77(1):43–63, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1859865.

Kristensen:2022:INWb

[KK22a] Hans M. Kristensen and Matt Korda. Indian nuclear weapons, 2022. Bulletin of the Atomic Scientists, 78(4):224–236, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2087385.

Kristensen:2022:INWa

[KK22b] Hans M. Kristensen and Matt Korda. Israeli nuclear weapons, 2021. Bulletin of the Atomic Scientists, 78(1):38–50, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.2014239.

Kristensen:2022:NKN

[KK22c] Hans M. Kristensen and Matt Korda. North Korean nuclear weapons, 2022. Bulletin of the Atomic Scientists, 78(5):273–294, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2109341.

Kristensen:2022:RNW

[KK22d] Hans M. Kristensen and Matt Korda. Russian nuclear weapons, 2022. Bulletin of the Atomic Scientists, 78(2):98–121, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2038907.

Kristensen:2022:USNa

[KK22e] Hans M. Kristensen and Matt Korda. United States nuclear weapons, 2022. Bulletin of the Atomic Scientists, 78(3):162–184, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2062943.

Kristensen:2022:USNb

[KK22f] Hans M. Kristensen and Matt Korda. United States nuclear weapons, 2023. Bulletin of the Atomic Scientists, 79(1):28–52, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2156686.

Kasperson:1979:IRT

[KKHK79] J. X. Kasperson, R. E. Kasperson, C. Hohenemser, and R. W. Kates. Institutional responses to Three Mile Island. Bulletin of the Atomic Scientists, 35(10):20–24, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sox80].

Kristensen:2023:FNW

[KKJ23a] Hans M. Kristensen, Matt Korda, and Eliana Johns. French nuclear weapons, 2023. Bulletin of the Atomic Scientists, 79 (4):272–281, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2223088.

Kristensen:2023:PNW

[KKJ23b] Hans M. Kristensen, Matt Korda, and Eliana Johns. Pakistan nuclear weapons, 2023. Bulletin of the Atomic Scientists, 79 (5):329–345, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2245260.

Kristensen:2023:CNWb

[KKJK23a] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. Chinese nuclear weapons, 2024. Bulletin of the Atomic Scientists, 80(1):49–72, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2295206.

Kristensen:2023:NWS

[KKJK23b] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. Nuclear weapons sharing, 2023. Bulletin of the Atomic Scientists, 79(6):393–406, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2266944.

Kristensen:2024:INW

[KKJK24a] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. Indian nuclear weapons, 2024. Bulletin of the Atomic Scientists, 80(5):326–342, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2388470.

Kristensen:2024:NKN

[KKJK24b] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. North Korean nuclear weapons, 2024. Bulletin of the Atomic Scientists, 80(4):251–271, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2365013.

Kristensen:2024:RNW

[KKJK24c] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. Russian nuclear weapons, 2024. Bulletin of the Atomic Scientists, 80(2):118–145, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2024.
2314437.

Kristensen:2024:UKN

[KKJK24d] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. United Kingdom nuclear weapons, 2024. Bulletin of the Atomic Scientists, 80(6):394–407, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2420550.

Kristensen:2024:USN

[KKJK24e] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. United States nuclear weapons, 2025. Bulletin of the Atomic Scientists, 81(1):53–79, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2024.
2441624.

Kristensen:2025:CNW

[KKJK25] Hans M. Kristensen, Matt Korda, Eliana Johns, and Mackenzie Knight. Chinese nuclear weapons, 2025. Bulletin of the Atomic Scientists, 81(2):135–160, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2025.
2467011.

Kang:2015:HPS

[KKL15] Jungmin Kang, Seoc-Woo Kim, and Byong-Chul Lee. Hot potato in South Korea: The spent nuclear fuel storage dilemma. Bulletin of the Atomic Scientists, 71(3):76–83, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2023:CNWa

[KKR23a] Hans M. Kristensen, Matt Korda, and Eliana Reynolds. Chinese nuclear weapons, 2023. Bulletin of the Atomic Scientists, 79 (2):108–133, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2178713.

Kristensen:2023:RNW

[KKR23b] Hans M. Kristensen, Matt Korda, and Eliana Reynolds. Russian nuclear weapons, 2023. Bulletin of the Atomic Scientists, 79 (3):174–199, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2202542.

Kutt:2022:RMV

[KKS22] Moritz Kütt, Ulrich Kühn, and Dmitry Stefanovich. Remote monitoring: Verifying geographical arms limits. Bulletin of the Atomic Scientists, 79(1):17–21, 2022. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2022.
2155006.

Kahin:1965:USV

[KL65] George McT. Kahin and John W. Lewis. The United States in Vietnam. Bulletin of the Atomic Scientists, 21(6):28–40, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kahin:1967:EEA

[KL67] George McT. Kahin and John W. Lewis. Escalation and East Asia. Bulletin of the Atomic Scientists, 23(1):20–24, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1976:PEA

[Kla76] Michael T. Klare. The political economy of arms sales. Bulletin of the Atomic Scientists, 32(9):10–18, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1977:ATA

[Kla77] Michael T. Klare. America’s top arms merchants. Bulletin of the Atomic Scientists, 33(6):20–21, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1979:IRT

[Kla79] Michael T. Klare. The international repression trade. Bulletin of the Atomic Scientists, 35(9):22–27, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1982:WYC

[Kla82] Michael T. Klare. The worst is yet to come. Bulletin of the Atomic Scientists, 38(6):45–46, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1984:SAT

[Kla84] Michael T. Klare. Soviet arms transfers to the Third World. Bulletin of the Atomic Scientists, 40(5):26–30, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1985:ONM

[Kla85] Michael T. Klare. 1985 outlook: New merchants in the arms bazaar. Bulletin of the Atomic Scientists, 41(1):15–17, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1988:SOC

[Kla88] Michael T. Klare. Secret operatives, clandestine trades: the thriving black market for weapons. Bulletin of the Atomic Scientists, 44(3):16–24, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1990:WGF

[Kla90] Michael T. Klare. Wars in the 1990s: Growing firepower in the Third World. Bulletin of the Atomic Scientists, 46(4):9–13, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1991:FFH

[Kla91] Michael T. Klare. Fueling the fire: how we armed the Middle East. Bulletin of the Atomic Scientists, 47(1):18–26, January/ February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1995:FV

[Kla95] Michael Klare. Flawed, but vital. Bulletin of the Atomic Scientists, 51(2):62, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1997:EAA

[Kla97a] Michael Klare. East Asia’s arms races. Bulletin of the Atomic Scientists, 53(1):18–19, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1997:EAM

[Kla97b] Michael Klare. East Asia’s militaries muscle up. Bulletin of the Atomic Scientists, 53(1):56–61, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1999:KA

[Kla99] Michael Klare. The Kalashnikov Age. Bulletin of the Atomic Scientists, 55(1):18–22, January/February 1999. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/18.

Klaus:2019:WRW

[Kla19] David M. Klaus. What really went wrong at WIPP: An insider’s view of two accidents at the only US underground nuclear waste repository. Bulletin of the Atomic Scientists, 75 (4):197–204, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1628516.

Klein:1981:CNR

[Kle81] John B. Klein. Commentary: No reason to worry. Bulletin of the Atomic Scientists, 37(9):64, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Con81b].

Kleinman:1991:LGW

[Kle91] Harold Kleinman. Letter: Gulf War a disaster. Bulletin of the Atomic Scientists, 47(5):46, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack91a].

Kline:1952:SSR

[Kli52] J. R. Kline. A symposium: The state of Russian science today: II: Soviet mathematics. Bulletin of the Atomic Scientists, 8(2): 44–47, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klingsberg:1959:LES

[Kli59] Erwin Klingsberg. Letter to the Editor: Soothing the savage beast? Bulletin of the Atomic Scientists, 15(9):401, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim59b].

Klipstein:1976:LRS

[Kli76] Pessa Klipstein. Letter: Radiological safety. Bulletin of the Atomic Scientists, 32(7):5, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klitgaard:1980:SYI

[Kli80] Robert E. Klitgaard. Support for young investigators. Bulletin of the Atomic Scientists, 36(10):20–23, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klima:1981:BRB

[Kli81] Otto Klima. Book review: Biological Paths to Self-Reliance: A Guide to Biological Solar Energy Conversion, by Russell E. Anderson. Bulletin of the Atomic Scientists, 37(7):46–47, August/ September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klintworth:1990:BRA

[Kli90] Gary Klintworth. Book review: Asian rivals: China eyes Japan, by Allen S. Whiting. Bulletin of the Atomic Scientists, 46(7):43– 44, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klose:1985:BRB

[Klo85] Kevin Klose. Book reviews: Behind the Lines: The Private War Against Soviet Censorship, by Donald R. Shanor; Whistleblowing in the Soviet Union: Complaints and Abuses under State Socialism. Bulletin of the Atomic Scientists, 41(8):54–55, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [KK85].

Klose:1986:RMF

[Klo86] Kevin Klose. Reporting from Moscow: fiction and secrets. Bulletin of the Atomic Scientists, 42(10):22–23, December 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kluge:2022:REM

[Klu22] Janis Kluge. Russia’s economy is much more than a “big gas station.” Under sanctions, that’s now its biggest problem. Bulletin of the Atomic Scientists, 78(6):307–309, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2132735.

Kennedy:1989:CSP

[KM89] William A. Kennedy and Mark G. Marshall. Commercial satellites: A peek at the French missile complex. Bulletin of the Atomic Scientists, 45(7):20–23, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Karns:1990:PES

[KM90] Margaret P. Karns and Karen A. Mingst. Peacekeeping efforts: Some fly, some flop. Bulletin of the Atomic Scientists, 46(4):

43–47, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krall:2018:BWP

[KM18] Lindsay Krall and Allison Macfarlane. Burning waste or playing with fire? Waste management considerations for nontraditional reactors. Bulletin of the Atomic Scientists, 74(5): 326–334, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1507791.

Kristensen:2004:PP

[KMN04] Hans M. Kristensen, Matthew G. McKinzie, and Robert S. Norris. The protection paradox. Bulletin of the Atomic Scientists, 60(2):68–77, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:/

/bos.sagepub.com/content/60/2/68;   http://docs.nrdc.
org/nuclear/nuc_04030101a_007.pdf.

Kristensen:2011:BNF

[KN11a] Hans M. Kristensen and Robert S. Norris. British nuclear forces, 2011. Bulletin of the Atomic Scientists, 67(5):89–97, September/ October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
5/89.

Kristensen:2011:CNF

[KN11b] Hans M. Kristensen and Robert S. Norris. Chinese nuclear forces, 2011. Bulletin of the Atomic Scientists, 67(6):81–87, November/ December 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
6/81.

Kristensen:2011:PNF

[KN11c] Hans M. Kristensen and Robert S. Norris. Pakistan’s nuclear forces, 2011. Bulletin of the Atomic Scientists, 67(4):91–99, July/ August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
4/91.

Kristensen:2011:RNF

[KN11d] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2011. Bulletin of the Atomic Scientists, 67(3):67–74, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/3/67.

Kristensen:2011:UNF

[KN11e] Hans M. Kristensen and Robert S. Norris. US nuclear forces, 2011. Bulletin of the Atomic Scientists, 67(2):66–76, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/2/66.

Kristensen:2012:INF

[KN12a] Hans M. Kristensen and Robert S. Norris. Indian nuclear forces, 2012. Bulletin of the Atomic Scientists, 68(4):96–101, July/ August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/68/
4/96.

Kristensen:2012:NNW

[KN12b] Hans M. Kristensen and Robert S. Norris. Nonstrategic nuclear weapons, 2012. Bulletin of the Atomic Scientists, 68(5):96–104, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/5/96.

Kristensen:2012:RNF

[KN12c] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2012. Bulletin of the Atomic Scientists, 68(2):87–97, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/2/87.

Kristensen:2012:UNF

[KN12d] Hans M. Kristensen and Robert S. Norris. US nuclear forces, 2012. Bulletin of the Atomic Scientists, 68(3):84–91, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/3/84.

Kristensen:2013:CNF

[KN13a] Hans M. Kristensen and Robert S. Norris. Chinese nuclear forces, 2013. Bulletin of the Atomic Scientists, 69(6):79–85, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/6/79.

Kristensen:2013:GNW

[KN13b] Hans M. Kristensen and Robert S. Norris. Global nuclear weapons inventories, 1945–2013. Bulletin of the Atomic Scientists, 69(5):75–81, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/5/75.

Kristensen:2013:RNF

[KN13c] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2013. Bulletin of the Atomic Scientists, 69(3):71–81, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/3/71.

Kristensen:2013:UNF

[KN13d] Hans M. Kristensen and Robert S. Norris. US nuclear forces, 2013. Bulletin of the Atomic Scientists, 69(2):77–86, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/2/77.

Kristensen:2014:BFN

[KN14a] Hans M. Kristensen and Robert S. Norris. The B61 family of nuclear bombs. Bulletin of the Atomic Scientists, 70(3):79–84, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2014:INW

[KN14b] Hans M. Kristensen and Robert S. Norris. Israeli nuclear weapons, 2014. Bulletin of the Atomic Scientists, 70(6):97–115, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2014:RNF

[KN14c] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2014. Bulletin of the Atomic Scientists, 70(2):75–85, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See corrigendum [Ano14a].

Kristensen:2014:SNW

[KN14d] Hans M. Kristensen and Robert S. Norris. Slowing nuclear weapon reductions and endless nuclear weapon modernizations: a challenge to the NPT. Bulletin of the Atomic Scientists, 70(4):

94–107, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2014:UNF

[KN14e] Hans M. Kristensen and Robert S. Norris. US nuclear forces, 2014. Bulletin of the Atomic Scientists, 70(1):85–93, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/70/1/
85.full.pdf+html.

Kristensen:2014:WDN

[KN14f] Hans M. Kristensen and Robert S. Norris. Worldwide deployments of nuclear weapons, 2014. Bulletin of the Atomic Scientists, 70(5):96–108, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2015:CNF

[KN15a] Hans M. Kristensen and Robert S. Norris. Chinese nuclear forces, 2015. Bulletin of the Atomic Scientists, 71(4):77–84, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2015:INF

[KN15b] Hans M. Kristensen and Robert S. Norris. Indian nuclear forces, 2015. Bulletin of the Atomic Scientists, 71(5):77–83, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/71/
5/77.full.pdf+html.

Kristensen:2015:PNF

[KN15c] Hans M. Kristensen and Robert S. Norris. Pakistani nuclear forces, 2015. Bulletin of the Atomic Scientists, 71(6):59–66, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2015:RNF

[KN15d] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2015. Bulletin of the Atomic Scientists, 71(3):84–97, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2015:UNF

[KN15e] Hans M. Kristensen and Robert S. Norris. US nuclear forces, 2015. Bulletin of the Atomic Scientists, 71(2):107–119, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2016:CNF

[KN16a] Hans M. Kristensen and Robert S. Norris. Chinese nuclear forces, 2016. Bulletin of the Atomic Scientists, 72(4):205–??, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2016:PNF

[KN16b] Hans M. Kristensen and Robert S. Norris. Pakistani nuclear forces, 2016. Bulletin of the Atomic Scientists, 72(6):368–376, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2016:RNF

[KN16c] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2016. Bulletin of the Atomic Scientists, 72(3):125–134, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2016:USN

[KN16d] Hans M. Kristensen and Robert S. Norris. United States nuclear forces, 2016. Bulletin of the Atomic Scientists, 72(2):63–73, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2017:HUN

[KN17a] Hans M. Kristensen and Robert S. Norris. A history of US nuclear weapons in South Korea. Bulletin of the Atomic Scientists, 73(6):349–357, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1388656.

Kristensen:2017:INF

[KN17b] Hans M. Kristensen and Robert S. Norris. Indian nuclear forces, 2017. Bulletin of the Atomic Scientists, 73(4):205–209, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1337998.

Kristensen:2017:RNF

[KN17c] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2017. Bulletin of the Atomic Scientists, 73(2):115–126, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1290375.

Kristensen:2017:USN

[KN17d] Hans M. Kristensen and Robert S. Norris. United States nuclear forces, 2017. Bulletin of the Atomic Scientists, 73(1):48– 57, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2016.1264213.

Kristensen:2017:WDN

[KN17e] Hans M. Kristensen and Robert S. Norris. Worldwide deployments of nuclear weapons, 2017. Bulletin of the Atomic Scientists, 73(5):289–297, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1363995.

Kristensen:2018:CNF

[KN18a] Hans M. Kristensen and Robert S. Norris. Chinese nuclear forces, 2018. Bulletin of the Atomic Scientists, 74(4):289–295, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1486620.

Kristensen:2018:NKN

[KN18b] Hans M. Kristensen and Robert S. Norris. North Korean nuclear capabilities, 2018. Bulletin of the Atomic Scientists, 74 (1):41–51, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1413062.

Kristensen:2018:RNF

[KN18c] Hans M. Kristensen and Robert S. Norris. Russian nuclear forces, 2018. Bulletin of the Atomic Scientists, 74(3):185–195, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1462912.

Kristensen:2018:USN

[KN18d] Hans M. Kristensen and Robert S. Norris. United States nuclear forces, 2018. Bulletin of the Atomic Scientists, 74(2): 120–131, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1438219.

Knapp:1954:SWL

[Kna54] Harold A. Knapp, Jr. South Woodley looks at the H-bomb. Bulletin of the Atomic Scientists, 10(8):306–311, October 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knapp:1963:PCD

[Kna63] Harold A. Knapp, Jr. Planning for civil defense: Five requirements. Bulletin of the Atomic Scientists, 19(4):39–41, 43, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelle:2010:SBP

[KND10] Alexander Kelle, Kathryn Nixdorff, and Malcolm Dando. Strengthening BWC prevention of state-sponsored bioweapons. Bulletin of the Atomic Scientists, 66(1):18–23, January 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/1/18.

Kristensen:2018:PNF

[KND18] Hans M. Kristensen, Robert S. Norris, and Julia Diamond. Pakistani nuclear forces, 2018. Bulletin of the Atomic Scientists, 74 (5):348–358, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1507796.

Kneese:1965:MHH

[Kne65] Allen V. Kneese. Man and his habitat: Problems of water pollution: New directions in water management. Bulletin of the Atomic Scientists, 21(5):2–8, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knorr:1955:DAW

[Kno55a] Klaus Knorr. Defense for atomic war. Bulletin of the Atomic Scientists, 11(3):77–81, 84, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knowland:1955:ASC

[Kno55b] Senator William Fife Knowland. Atomic stalemate and coexistence. Bulletin of the Atomic Scientists, 11(1):28, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knorr:1960:FWD

[Kno60] Klaus Knorr. The future of Western deterrent power. II. A view from the United States. Bulletin of the Atomic Scientists, 16(7): 271–276, 282, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knorr:1966:CEA

[Kno66] Klaus Knorr. On the cost-effectiveness approach to military research and development. Bulletin of the Atomic Scientists, 22 (9):11–14, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Knoll:1990:LHB

[Kno90] Erwin Knoll. Letter: H-bomb secret never was. Bulletin of the Atomic Scientists, 46(3):53–54, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90a].

Knudson:1989:RRR

[Knu89] Tom Knudson. Report: Risk reduction: Rancho Seco decked again. Bulletin of the Atomic Scientists, 45(10):7–8, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Koblick:1970:TTI

[Kob70] Ian Koblick. Trends: Tektite III, anyone? Bulletin of the Atomic Scientists, 26(8):30, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Koch:1957:LEC

[Koc57] Leo F. Koch. Letter to the Editor: A Christian answer to war. Bulletin of the Atomic Scientists, 13(1):40–41, January 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [McC56, Dav57a].

Kochs:1977:LGV

[Koc77] Dorothy Kochs. Letter: Genetic vulnerability. Bulletin of the Atomic Scientists, 33(5):2, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil77a].

Koch:1998:LWI

[Koc98] Andrew R. Koch. Letter: Work with Iran. Bulletin of the Atomic Scientists, 54(2):69, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Arn98c].

Koch:2004:DBD

[Koc04] Lewis Z. Koch. Dirty bomber? dirty justice. Bulletin of the Atomic Scientists, 60(1):59–68, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/1/59.

Koenig:1977:TTW

[Koe77] M. E. D. Koenig. The toy theory of Western history. Bulletin of the Atomic Scientists, 33(8):16–18, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Wie78, Fri78, Bar78c].

Koestler:1979:PJS

[Koe79] Arthur Koestler. Perspective: JANUS: a summing up. Bulletin of the Atomic Scientists, 35(3):4, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from the author’s 1978 book JANUS. See comment [Coo79].

Konovalov:1993:SSS

[KOE93] Alexander Konovalov, Sergei Oznobistchev, and Dmitri Evstafiev. Sanctions: Saying da, saying nyet. Bulletin of the Atomic Scientists, 49(9):28–31, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Koester:2022:BBD

[Koe22] Stefan Koester. Burning biomass: A Drax-tic idea, and bad for environmental justice. Bulletin of the Atomic Scientists, 78 (3):148–151, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2062939.

Kohahl:1976:LBB

[Koh76] Robert E. Kohahl. Letter: Body and the Bible. Bulletin of the Atomic Scientists, 32(2):4, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pre75].

Kokoshin:1988:RFE

[Kok88a] Andrei A. Kokoshin. Restructure forces, enhance security. Bulletin of the Atomic Scientists, 44(7):35–38, September 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kokoshin:1988:SVR

[Kok88b] Andrei A. Kokoshin. A Soviet view on radical weapons cuts. Bulletin of the Atomic Scientists, 44(2):14–17, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kolkowicz:1963:BBC

[Kol63] Roman Kolkowicz. Books: The Challenge of Coexistence: A Study of Soviet Economic Diplomacy, by Milton Kovner. Bulletin of the Atomic Scientists, 19(3):26–27, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kolodziej:1964:FWD

[Kol64] Edward A. Kolodziej. Forum: What does the test ban mean? Bulletin of the Atomic Scientists, 20(3):22–23, March 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kolmas:1966:MN

[Kol66] Josef Kolmas. The minority nationalities. Bulletin of the Atomic Scientists, 22(6):71–74, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kolodiy:1975:VPC

[Kol75] George Kolodiy. Velikovsky: Paradigms in collision: The challenging of established scientific beliefs. Bulletin of the Atomic Scientists, 31(2):36–38, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dun75].

Kolodziej:1981:CFL

[Kol81] Edward A. Kolodziej. Commentary: Furthermore …. Bulletin of the Atomic Scientists, 37(7):25, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gal81b].

Kolodziej:1984:CFD

[Kol84] Edward A. Kolodziej. Commentary: On ‘The French defense’. Bulletin of the Atomic Scientists, 40(6):50, June/July 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [FC84].

Kolb:2007:UTE

[Kol07] Rocky Kolb. A unified theory of Einstein. Bulletin of the Atomic Scientists, 63(6):58–59, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/63/6/58;
 http://bos.sagepub.com/content/63/6/58.abstract. Re-

view of [Isa07].

Komai:1954:SFO

[Kom54] Taku Komai. Science and freedom in the Orient. Bulletin of the Atomic Scientists, 10(6):209–210, 214, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Komanoff:1980:CUN

[Kom80] Charles Komanoff. Commentary: U.S. nuclear plant performance. Bulletin of the Atomic Scientists, 36(9):51–54, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Komisar:1990:REE

[Kom90] Lucy Komisar. Report: East erupts, NATO squirms. Bulletin of the Atomic Scientists, 46(1):7–9, January/February 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kontos:1967:BBF

[Kon67] C. William Kontos. Books: Planning Without Facts: Lessons in Resource Allocation From Nigeria’s Development, by Wolfgang F. Stolper. Bulletin of the Atomic Scientists, 23(6):53–54, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kopytko:2015:USU

[Kop15] Natalie Kopytko. Uncertain seas, uncertain future for nuclear power. Bulletin of the Atomic Scientists, 71(2):29–38, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Korf:1964:LEF

[Kor64] Stanley R. Korf. Letter to the Editor: Fluoridation. Bulletin of the Atomic Scientists, 20(7):30, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar64h].

Korman:1972:CPP

[Kor72] Henry R. Korman. Communications: Power policy. Bulletin of the Atomic Scientists, 28(10):2, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil72b].

Korman:1975:LBW

[Kor75] Henry R. Korman. Letters: A better way. Bulletin of the Atomic Scientists, 31(3):2, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [YY74].

Kornetchuk:1977:PSA

[Kor77] Elena Kornetchuk. The politics of Soviet art. Bulletin of the Atomic Scientists, 33(8):32–37, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Korn:1989:LKV

[Kor89] David A. Korn. Letter: The Kremlin vote. Bulletin of the Atomic Scientists, 45(2):49, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gad89].

Kortunov:1990:SIB

[Kor90] Sergei Kortunov. START II and beyond. Bulletin of the Atomic Scientists, 46(8):21–24, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kornbluh:1992:BRG

[Kor92] Mark L. Kornbluh. Book review: Giving aid and comfort: Secret Agenda: The United States Government, Nazi Scientists and Project Paperclip, 1945–1990, by Linda Hunt. Bulletin of the Atomic Scientists, 48(7):42–43, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Korb:2010:IQH

[Kor10] Lawrence J. Korb. The Iranian quagmire: How to move forward: Position: Containment and deterrence. Bulletin of the Atomic Scientists, 66(6):95–101, November/December 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/6/95.

Korzak:2016:QCN

[Kor16] Elaine Korzak. The quest for cyber norms. Bulletin of the Atomic Scientists, 72(5):348–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Korb:2017:RTN

[Kor17] Lawrence J. Korb. Rising tensions, nuclear modernizations: How Washington can turn down the heat. Bulletin of the Atomic Scientists, 73(3):173–176, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1315038.

Korb:2019:FUM

[Kor19] Lawrence J. Korb. The focus of US military efforts in outer space should be … arms control. Bulletin of the Atomic Scientists, 75(4):148–150, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1628471.

Korb:2021:WPB

[Kor21] Lawrence Korb. Why President Biden needs to revisit — and reduce — his defense budget. Bulletin of the Atomic Scientists, 77(5):249–250, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1964234.

Kosower:1960:LEO

[Kos60] Edward M. Kosower. Letter to the Editor: The Only hope? Bulletin of the Atomic Scientists, 16(1):0, 33, January 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Doo59].

Kosal:2004:SS

[Kos04a] Margaret E. Kosal. Is small scary? Bulletin of the Atomic Scientists, 60(5):38–47, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/5/38.

Koster:2004:BD

[Kos04b] Karel Koster. The best defense. … Bulletin of the Atomic Scientists, 60(5):26–28, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/5/26.

Kosal:2010:SIN

[Kos10] Margaret Kosal. The security implications of nanotechnology. Bulletin of the Atomic Scientists, 66(4):58–69, July/August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/4/58.

Kosal:2019:TN

[Kos19] Margaret E. Kosal. The threats from nanotechnology. Bulletin of the Atomic Scientists, 75(6):290–294, 2019. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1680054.

Kouts:1975:NSR

[Kou75] H. J. C. Kouts. NRC safety research which may be expanded. Bulletin of the Atomic Scientists, 31(7):34–35, September 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kourous:2001:P

[Kou01] George Kourous. On patrol. Bulletin of the Atomic Scientists, 57(2):73–74, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/2/73.

Kovda:1968:SUS

[Kov68] Victor A. Kovda. Search for a U.N. science policy. Bulletin of the Atomic Scientists, 24(3):12–16, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kowarski:1948:AED

[Kow48] L. Kowarski. Atomic energy developments in France. Bulletin of the Atomic Scientists, 4(5):139–140, 154–155, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kowarski:1949:PSL

[Kow49] L. Kowarski. Psychology and structure of large-scale physical research. Bulletin of the Atomic Scientists, 5(6–7):186–191, 200, 204, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kowarski:1955:MCE

[Kow55] L. Kowarski. The making of CERN — an experiment in cooperation. Bulletin of the Atomic Scientists, 11(10):354–357, 381, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kowarski:1977:CSI

[Kow77] Lew Kowarski. Conditions of success in international enterprises in science and technology. Bulletin of the Atomic Scientists, 33 (7):45–48, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristoferson:1971:CRS

[KP71] Lars Kristoferson and Victor Paschkis. Communications: Relevance of space. Bulletin of the Atomic Scientists, 27(10):2–3, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kan71a, Kan71b].

Krepon:1985:AAS

[KP85] Michael Krepon and D. Geoffrey Peck. Another alarm on Soviet ABMs. Bulletin of the Atomic Scientists, 41(6):34–36, June/ July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Koch:1998:SM

[KP98] Andrew Koch and Waheguru Pal Singh Sidhu. Subcontinental missiles. Bulletin of the Atomic Scientists, 54(4):44–49, July/ August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kincade:1984:CNT

[KR84] William H. Kincade and Carol V. Rose. Commentary: No technical fix. Bulletin of the Atomic Scientists, 40(6):48, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kelleher:2011:GZP

[KR11] Catherine McArdle Kelleher and Judith (Judith Voris) Reppy, editors. Getting to zero: the path to nuclear disarmament. Stanford Security Studies, Stanford, CA, USA, 2011. ISBN 0-8047-7394-7. xx + 404 pp. LCCN JZ5675 .G49 2011.

Korb:2012:NFU

[KR12] Lawrence J. Korb and Alexander Rothman. No first use: The way to contain nuclear war in South Asia. Bulletin of the Atomic Scientists, 68(2):34–42, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/2/34.

Kramers:1946:SRI

[Kra46] H. A. Kramers. The scientists’ role in international relations. Bulletin of the Atomic Scientists, 2(11–12):12–13, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kramish:1957:LEA

[Kra57] Arnold Kramish. Letter to the Editor: After Pugwash: The Soviet reaction. Bulletin of the Atomic Scientists, 13(10):377, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab57c, Hoc57].

Kramish:1959:AEU

[Kra59] Arnold Kramish. Atomic energy in the USSR. Bulletin of the Atomic Scientists, 15(8):322–328, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krause:1972:CDO

[Kra72] Kurt W. Krause. Communications: Difference of opinion. Bulletin of the Atomic Scientists, 28(7):3, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Str72, Haa73].

Kramer:1973:ECW

[Kra73] Eugene Kramer. Energy conservation and waste recycling: Taking advantage of urban congestion. Bulletin of the Atomic Scientists, 29(4):13–18, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kramish:1978:SRM

[Kra78] Arnold Kramish. Smyth report mystery. Bulletin of the Atomic Scientists, 34(6):5, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kramer:1989:LSR

[Kra89] Mark Kramer. Letter: Skeptic replies. Bulletin of the Atomic Scientists, 45(8):43–44, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mye90].

Krass:1991:SPD

[Kra91] Allan S. Krass. START: The people, the debt, and Mikhail. Bulletin of the Atomic Scientists, 47(9):12–17, November 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krasno:1992:RNB

[Kra92] Jean Krasno. Report: Nonproliferation: Brazil, Argentina make it official. Bulletin of the Atomic Scientists, 48(3):10–11, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krass:1994:BRF

[Kra94] Allan S. Krass. Book review: Flawed theory: We All Lost the Cold War, by Richard Ned Lebow and Janice Gross Stein. Bulletin of the Atomic Scientists, 50(5):54–56, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kraig:1999:SSY

[Kra99] Michael Kraig. Safe or sorry: The “Y2K Problem” and nuclear weapons. Bulletin of the Atomic Scientists, 55(2):38–43, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/2/38.

Krauss:2006:EJS

[Kra06] Lawrence M. Krauss. Embracing junk science. Bulletin of the Atomic Scientists, 62(6):51–53, November/December 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/6/51.

Krauss:2007:CDR

[Kra07] Lawrence M. Krauss. A case of dubious rationales. Bulletin of the Atomic Scientists, 63(4):41–42, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/41.

Krauss:2010:CMDa

[Kra10] Lawrence M. Krauss. Critical mass: The Doomsday Clock still ticks. Scientific American, 302(1):40, January 2010. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v302/n1/full/scientificamerican0110-40.html;   http:/
/www.nature.com/scientificamerican/journal/v302/n1/pdf/
scientificamerican0110-40.pdf.

Krauss:2015:SCB

[Kra15] Lawrence M. Krauss. Scientists as celebrities: Bad for science or good for society? Bulletin of the Atomic Scientists, 71(1): 26–32, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krenz:1977:LEV

[Kre77] Jerrold H. Krenz. Letter: Efficiency vs. activity. Bulletin of the Atomic Scientists, 33(2):2–3, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RW76].

Krenz:1981:NED

[Kre81] Jerrold H. Krenz. The nation’s energy dilemma. Bulletin of the Atomic Scientists, 37(5):31–33, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1983:BRB

[Kre83a] Michael Krepon. Book review: Disarming Europe, edited by Mary Kaldor and Dan Smith. Bulletin of the Atomic Scientists, 39(9):45–46, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1983:SH

[Kre83b] Michael Krepon. START on hold. Bulletin of the Atomic Scientists, 39(2):7–9, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1985:BRB

[Kre85a] Michael Krepon. Book review: Decisionmaking for Arms Limitation, edited by Hans Guenter Brauch and Duncan Clarke. Bulletin of the Atomic Scientists, 41(3):56–57, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1985:PTW

[Kre85b] Michael Krepon. Perspective: Technology won’t solve verification problems. Bulletin of the Atomic Scientists, 41(2):3–4, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1986:DTA

[Kre86a] Michael Krepon. Dormant threat to the ABM Treaty. Bulletin of the Atomic Scientists, 42(1):31–34, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1986:WRM

[Kre86b] Michael Krepon. Washington report: Mixed signals on arms control. Bulletin of the Atomic Scientists, 42(7):6–7, August/ September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1986:WRN

[Kre86c] Michael Krepon. Washington report: Neoconservative war of the worlds. Bulletin of the Atomic Scientists, 42(3):6–7, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1986:WRR

[Kre86d] Michael Krepon. Washington report: Ronald Reagan’s hidden hands. Bulletin of the Atomic Scientists, 42(5):4–5, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1986:WRP

[Kre86e] Michael Krepon. Washington report: The PR administration. Bulletin of the Atomic Scientists, 42(9):6–7, November 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1987:HSI

[Kre87a] Michael Krepon. High stakes in INF verification. Bulletin of the Atomic Scientists, 43(5):14–16, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1987:WRC

[Kre87b] Michael Krepon. Washington report: CIA, DIA at odds over Soviet threat. Bulletin of the Atomic Scientists, 43(4):6–7, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1987:WRIb

[Kre87c] Michael Krepon. Washington report: INF agreement in principle. Bulletin of the Atomic Scientists, 43(9):5–6, November 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1987:WRR

[Kre87d] Michael Krepon. Washington report: Rocky road to INF accord. Bulletin of the Atomic Scientists, 43(7):3–4, September 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1987:WRIa

[Kre87e] Michael Krepon. Washington report: The Iran/arms control connection. Bulletin of the Atomic Scientists, 43(2):9–10, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1987:WRS

[Kre87f] Michael Krepon. Washington report: The surprise defense initiative. Bulletin of the Atomic Scientists, 43(6):5–6, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kreidler:1988:CRP

[Kre88a] William J. Kreidler. Conflict resolution on the planet Zimock. Bulletin of the Atomic Scientists, 44(6):20, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1988:CAS

[Kre88b] Michael Krepon. Can this agency be saved? Bulletin of the Atomic Scientists, 44(10):35–38, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wal89b].

Krepon:1988:PCY

[Kre88c] Michael Krepon. The presidential campaign: You can run and you can hide. Bulletin of the Atomic Scientists, 44(9):6–7, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1988:WRA

[Kre88d] Michael Krepon. Washington report: Arms control play-by-play. Bulletin of the Atomic Scientists, 44(2):5–6, March 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1988:WRN

[Kre88e] Michael Krepon. Washington report: Nunn’s modest SDI. Bulletin of the Atomic Scientists, 44(3):5, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1988:WRS

[Kre88f] Michael Krepon. Washington report: STARTing a new Soviet menace? Bulletin of the Atomic Scientists, 44(6):5–6, July/ August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tsi89b].

Krepon:1989:ACC

[Kre89a] Michael Krepon. Arms control criteria: Has arms control worked? Bulletin of the Atomic Scientists, 45(4):26–28, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1989:CSP

[Kre89b] Michael Krepon. Commercial satellites: Peacemakers or rent-a-spies? Bulletin of the Atomic Scientists, 45(7):12–15, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kretzmann:1989:BRW

[Kre89c] Mark Kretzmann. Book review: Where are they now? Unit 731: Japan’s Secret Biological Warfare in World War II, by Peter Williams and David Wallace. Bulletin of the Atomic Scientists, 45 (10):47–48, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1990:BRT

[Kre90] Michael Krepon. Book review: Two spins on arms control à la Reagan: The Great Universal Embrace: Arms Summitry, A Skeptic’s Account, by Kenneth L. Adelman; From Hiroshima to Glasnost: At the Center of Decision — A Memoir, by Paul H. Nitze with Ann M. Smith and Steven L. Reardon. Bulletin of the Atomic Scientists, 46(6):41–42, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1991:NWO

[Kre91a] Michael Krepon. New world order: An odor of the old. Bulletin of the Atomic Scientists, 47(5):37, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1991:PDP

[Kre91b] Michael Krepon. Perspective: Don’t parrot old arguments on missile defense. Bulletin of the Atomic Scientists, 47(1):12–13, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1992:PLO

[Kre92a] Michael Krepon. Perspective: Less offense, more defense. Bulletin of the Atomic Scientists, 48(3):12–13, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano92-30].

Krepon:1992:WWF

[Kre92b] Michael Krepon. Where will the “fool’s gold” go? Bulletin of the Atomic Scientists, 48(4):31–32, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1993:BRT

[Kre93] Michael Krepon. Book review: Too many cooks? Arms Control by Committee: Managing Negotiations with the Russians, by George Bunn. Bulletin of the Atomic Scientists, 49(9):51– 52, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1995:PDI

[Kre95] Michael Krepon. Perspective: Do it better and smarter. Bulletin of the Atomic Scientists, 51(6):13–14, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1996:CR

[Kre96] Michael Krepon. The counter-revolution. Bulletin of the Atomic Scientists, 52(5):45–46, September/October 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1997:GOI

[Kre97] Michael Krepon. Guest opinion: Isolate Helms; push treaties. Bulletin of the Atomic Scientists, 53(1):4, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kreuter:1998:BRF

[Kre98] Gretchen Kreuter. Book review: The forgotten holocaust: The Rape of Nanking: The Forgotten Holocaust of World War II, by Iris Chang. Bulletin of the Atomic Scientists, 54(2):65–66, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krepon:1999:TTC

[Kre99] Michael Krepon. Troubled treaties: CTBT deadline nears. Bulletin of the Atomic Scientists, 55(2):18–20, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/2/18.

Krepon:2003:DR

[Kre03] Michael Krepon. Dominators rule. Bulletin of the Atomic Scientists, 59(1):55–60, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/55.

Krepon:2005:GRS

[Kre05] Michael Krepon. Ground rules for space. Bulletin of the Atomic Scientists, 61(3):68, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/68.

Krepon:2007:NF

[Kre07] Michael Krepon. The nuclear flock. Bulletin of the Atomic Scientists, 63(2):15–17, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/2/15.

Krebs:2011:LCC

[Kre11] Martha Krebs. Lessons of climate change, stories of solutions: California: Government leadership. Bulletin of the Atomic Scientists, 67(1):60–63, January/February 2011. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/1/60.

Krieger:1975:TNT

[Kri75] David Krieger. Terrorists and nuclear technology: The danger is great; the question is not whether the worst will happen, but where and how. Bulletin of the Atomic Scientists, 31(6):28–34, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krimsky:1978:CCR

[Kri78] Sheldon Krimsky. A citizen court in the recombinant DNA debate. Bulletin of the Atomic Scientists, 34(8):37–43, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krimsky:1985:BRB

[Kri85] Sheldon Krimsky. Book reviews: No Fire, No Thunder, by Sean Murphy, Alastair Hay, and Steven Rose; Environmental Warfare: A Technical, Legal and Policy Appraisal, edited by Arthur H. Westing; Herbicides in War: The Long-Term Ecological and Human Consequences, edited by Arthur H. Westing. Bulletin of the Atomic Scientists, 41(8):55–57, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krinsky:1986:CPA

[Kri86] Robert Krinsky. Community programs aid Pentagon. Bulletin of the Atomic Scientists, 42(8):37–39, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krieger:1987:LNT

[Kri87] David Krieger. Letter: Nuclear terrorist states. Bulletin of the Atomic Scientists, 43(2):62, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Int86b].

Kriedman:1991:LSW

[Kri91] Herbert Kriedman. Letter: Say what you mean. Bulletin of the Atomic Scientists, 47(6):45, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap91a].

Kriedman:1992:LAD

[Kri92a] Herbert Kriedman. Letter: Arms diplomacy. Bulletin of the Atomic Scientists, 48(8):47, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:1992:BNC

[Kri92b] Hans M. Kristensen. Bulletin: Neither confront nor deny. Bulletin of the Atomic Scientists, 48(2):3–4, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kriedman:1995:LAS

[Kri95] Herbert Kriedman. Letter: Anti-Serb reporting. Bulletin of the Atomic Scientists, 51(5):60, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Meh95].

Kristensen:1997:TO

[Kri97] Hans Kristensen. Targets of opportunity. Bulletin of the Atomic Scientists, 53(5):22–28, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Roc97a, SC97].

Krieger:1998:LWA

[Kri98] David Krieger. Letter: We are all culpable. Bulletin of the Atomic Scientists, 54(4):3–4, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [But98].

Kristensen:2002:BDD

[Kri02a] Hans M. Kristensen. Bomb deal a dud, part one. Bulletin of the Atomic Scientists, 58(1):25–26, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/25.

Kristensen:2002:PP

[Kri02b] Hans M. Kristensen. Preemptive posturing. Bulletin of the Atomic Scientists, 58(5):54, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kristensen:2002:PPW

[Kri02c] Hans M. Kristensen. Preemptive posturing: What happened to deterrence? Bulletin of the Atomic Scientists, 58(5):54–59, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/5/54.

Krige:2014:NSA

[Kri14] John Krige. National security and academia: Regulating the international circulation of knowledge. Bulletin of the Atomic Scientists, 70(2):42–52, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kroenig:2015:HAN

[Kro15] Matthew Kroenig. How to approach nuclear modernization? : a US response. Bulletin of the Atomic Scientists, 71(3):16–18, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kruzel:1974:BBS

[Kru74] Joseph Kruzel. Books: The Superpowers and Arms Control, by Walter C. Clemens, Jr. Bulletin of the Atomic Scientists, 30(6): 58–60, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [CO74].

Krugmann:1981:GBN

[Kru81] Hartmut Krugmann. The German–Brazilian nuclear deal. Bulletin of the Atomic Scientists, 37(2):32–36, February 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Krzyzaniak:2020:IBB

[Krz20] John Krzyzaniak. Interview: Brian Brettschneider: How climate change has already arrived in the Arctic. Bulletin of the Atomic Scientists, 76(3):129–132, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1751964.

Kellog:1958:SNW

[KS58] W. W. Kellog and Charles Schafer. Supplement: Nuclear weapons in war: Fall-out from a bombing campaign. Bulletin of the Atomic Scientists, 14(1):59–61, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Khariton:1993:KV

[KS93] Yuli Khariton and Yuri Smirnov. The Khariton version. Bulletin of the Atomic Scientists, 49(4):20–31, May 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.thebulletin.org/article.php?
art_ofn=may93khariton. See comment [Hol94c].

Kaplan:1994:LZO

[KS94] Martin Kaplan and Lloyd G. Shore. Letters: Zero — the only option. Bulletin of the Atomic Scientists, 50(1):62, January/ February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo93b].

Khripunov:1998:RMR

[KS98] Igor Khripunov and Anupam Srivastava. From Russia, a muted reaction. Bulletin of the Atomic Scientists, 54(4):42–43, July/ August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ksanfomality:1991:SBS

[Ksa91] Leonid V. Ksanfomality. Survival before science. Bulletin of the Atomic Scientists, 47(10):20–24, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Korbatov:2015:FAN

[KSG15] Anna Bella Korbatov, Erika Suzuki, and Bethany L. Goldblum. The fight against nuclear terrorism needs global cooperation — and the IAEA. Bulletin of the Atomic Scientists, 71(5): 67–76, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/67.full.pdf+html.

Kuhn:2017:ENW

[KSS17] Ulrich Kühn, Shatabhisha Shetty, and Polina Sinovets. Europe’s nuclear woes: Mitigating the challenges of the next years. Bulletin of the Atomic Scientists, 73(4):245–254, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1338022.

Kellman:1997:CTD

[KT97] Barry Kellman and Edward A. Tanzman. Chemical treaty deserves ratification. Bulletin of the Atomic Scientists, 53(1):15– 17, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kubbig:1988:SWF

[Kub88] Bernd W. Kubbig. Star Wars fizzles for European contractors. Bulletin of the Atomic Scientists, 44(9):17–20, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kuenzli:1961:LER

[Kue61] Alfred E. Kuenzli. Letter to the Editor: Response to Wigner. Bulletin of the Atomic Scientists, 17(5–6):256, May/June 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wig61].

Kugler:1969:CGR

[Kug69] Israel Kugler. Communications: “Grave restiveness”. Bulletin of the Atomic Scientists, 25(10):43, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Din69].

Kuhn:1966:LEJ

[Kuh66] Philip Kuhn. Letter to the Editor: From Japan. Bulletin of the Atomic Scientists, 22(3):26–27, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kuhfeld:1975:LS

[Kuh75] Albert W. Kuhfeld. Letters: Not surprising. Bulletin of the Atomic Scientists, 31(3):2, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bur74].

Kuhn:2017:IND

[Küh17] Ulrich Kühn. Introduction: Nuclear disarmament and arms control for the next decade. Bulletin of the Atomic Scientists, 73(4):244, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1338020.

Kuhn:2018:DD

[Küh18] Ulrich Kühn. Deterrence and its discontents. Bulletin of the Atomic Scientists, 74(4):248–254, 2018. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2018.
1486613.

Kuiper:1973:APL

[Kui73] Gerard P. Kuiper. The Apollo Program and lunar science. Bulletin of the Atomic Scientists, 29(10):19–26, December 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kull:1988:FGA

[Kul88] Steven Kull. Feeling good about hard-target-kill capability. Bulletin of the Atomic Scientists, 44(6):30–35, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kull:1995:MPM

[Kul95] Steven Kull. Misreading the public mood. Bulletin of the Atomic Scientists, 51(2):55–59, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kulacki:2006:LT

[Kul06] Gregory Kulacki. Lost in translation. Bulletin of the Atomic Scientists, 62(3):34–39, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/3/34.

Kulacki:2017:CPR

[Kul17] Gregory Kulacki. China’s proper role in the global nuclear order. Bulletin of the Atomic Scientists, 73(2):131–132, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1288452.

Kumaraswamy:1995:RME

[Kum95] P. R. Kumaraswamy. Report: Middle East: Egypt needles Israel. Bulletin of the Atomic Scientists, 51(2):11–13, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kumaraswamy:1999:IES

[Kum99] P. R. Kumaraswamy. Israel: an embarrassment of spies. Bulletin of the Atomic Scientists, 55(2):??, March 1999. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/2/14.

Kumaraswamy:2004:ASI

[Kum04] P. R. Kumaraswamy. Arms sales: Are Indo–Israeli deals doomed? Bulletin of the Atomic Scientists, 60(6):12–13, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/6/12.

Kuntzel:1996:LGO

[Kün96] Matthias Küntzel. Letter: Germany’s other plutonium option. Bulletin of the Atomic Scientists, 52(4):3, July/August 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hib96b].

Kunihiro:1997:DFP

[Kun97] Masao Kunihiro. The decline and fall of pacifism. Bulletin of the Atomic Scientists, 53(1):35–39, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kunertova:2023:WUS

[Kun23] Dominika Kunertova. The war in Ukraine shows the gamechanging effect of drones depends on the game. Bulletin of the Atomic Scientists, 79(2):95–102, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2178180.

Kuper:1951:RMC

[Kup51] J. B. H. Kuper. Radiological monitoring for civil defense. Bulletin of the Atomic Scientists, 7(10):313–314, 320, October 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kuperman:2006:BGB

[Kup06] Alan J. Kuperman. Bomb-grade bazaar. Bulletin of the Atomic Scientists, 62(2):44–50, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/2/44.

Kurchatov:1956:CTR

[Kur56] Igor Kurchatov. Controlled thermonuclear reactions. Bulletin of the Atomic Scientists, 12(7):269–271, September 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Pravda, 10 May 1956.

Kurihara:1979:ADP

[Kur79] Sadako Kurihara. America, do not perish at your hands — in protest of the Texas Air Show. Bulletin of the Atomic Scientists, 35(1):14, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Poem in protest of reenactment of atomic bombing of Hiroshima, republished from Peace and Change: A Journal of Peace Research.

Kusch:1968:WSS

[Kus68] Polykarp Kusch. The world of science and the scientist’s world. Bulletin of the Atomic Scientists, 24(8):38–43, October 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kusnitz:1983:BRB

[Kus83] Leonard Kusnitz. Book review: The Dragon and the Bear: Inside China and Russia Today, by Philip Short. Bulletin of the Atomic Scientists, 39(7):51, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kuyper:1954:LNA

[Kuy54] Adrian Kuyper. A look at the new Atomic Energy Law. Bulletin of the Atomic Scientists, 10(10):389–392, December 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kuzin:1958:RMB

[Kuz58] A. M. Kuzin. Radiation and man: Broad studies: Some current problems in radiobiology. Bulletin of the Atomic Scientists, 14(1): 48–51, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Kuznick:1988:BSA

[Kuz88] Peter J. Kuznick. The birth of scientific activism. Bulletin of the Atomic Scientists, 44(10):39–40, 42–43, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mar89b].

Kuznick:2004:SS

[Kuz04] Peter J. Kuznick. Scientists on the stump. Bulletin of the Atomic Scientists, 60(6):28–35, November/December 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/6/28.

Kogut:1986:TPA

[KW86] John Kogut and Michael Weissman. Taking the pledge against Star Wars. Bulletin of the Atomic Scientists, 42(1):27–30, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bal86a].

Kaiser:2015:ASP

[KW15] David Kaiser and Benjamin Wilson. American scientists as public citizens: 70 years of the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 71(1):13–25, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Klare:1988:LND

[KWM88] Michael Klare, Jan I. Wolitzky, and Donald J. Montgomery. Letter: Nonoffensive defense. Bulletin of the Atomic Scientists, 44 (10):49–50, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vMK88].

Kwon:2010:NND

[Kwo10] Hee-Seog Kwon. Negotiating with the North: Doubting its enrichment claims. Bulletin of the Atomic Scientists, 66(3):10–18, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/66/3/10.

Kyl:1997:LCT

[Kyl97] Jon Kyl. Letter: Chemical treaty deeply flawed. Bulletin of the Atomic Scientists, 53(1):3, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa96d].

KCS:1985:JSH

[Kyo85] Kyoto Conference of Scientists. Japanese scientists hit military R&D. Bulletin of the Atomic Scientists, 41(2):43, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFave:1966:LEN

[La 66] Lawrence La Fave. Letter to the Editor: The need for a third party. Bulletin of the Atomic Scientists, 22(8):40, October 1966.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaRocque:1978:LSP

[La 78] Gene R. La Rocque. Letter: Strategic parity. Bulletin of the Atomic Scientists, 34(2):3, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tho77].

LaRocque:1979:SRF

[La 79] Gene R. La Rocque. Special report: The first nuclear war conference. Bulletin of the Atomic Scientists, 35(4):20, April 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFeber:1983:OCA

[La 83] Walter La Feber. An overview of Central America. Bulletin of the Atomic Scientists, 39(8):11–13, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFarge:1988:MGE

[La 88a] Phyllis La Farge. Mainstreaming global education. Bulletin of the Atomic Scientists, 44(6):16, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFarge:1988:NTP

[La 88b] Phyllis La Farge. Nuclear teaching: propaganda or problem solving? Bulletin of the Atomic Scientists, 44(6):14–20, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cam88].

LaFarge:1988:OE

[La 88c] Phyllis La Farge. The Oregon experiment. Bulletin of the Atomic Scientists, 44(6):18, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFeber:1991:BRL

[La 91] Walter La Feber. Book review: Losers’ history: The Vietnam Wars, 1945–1990, by Marilyn B. Young. Bulletin of the Atomic Scientists, 47(10):40–41, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lorber:1998:LCM

[LA98] Azriel Lorber and Eric Arnett. Letter: Chinese missiles. Bulletin of the Atomic Scientists, 54(1):68–69, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Arn97b].

LaBrant:1981:CLM

[LaB81] Lou LaBrant. Commentary: Like-minded writers and readers. Bulletin of the Atomic Scientists, 37(2):64, February 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lacassagne:1957:RCF

[Lac57] A. Lacassagne. The risks of cancer formation by radiations. Bulletin of the Atomic Scientists, 13(4):135–136, 142, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lacis:1969:UDA

[Lac69] Valdis Lim Lacis. Undersea defenses against China. Bulletin of the Atomic Scientists, 25(5):50, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lach:1985:BRB

[Lac85] Joseph Lach. Book reviews: Countdown to Space War, by Bhupendra Jasani and Christopher Lee; National Interests and the Military Use of Space, edited by William L. Durch; The Fallacy of Star Wars, edited by John Tirman. Bulletin of the Atomic Scientists, 41(6):50–51, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lach:1990:LBP

[Lac90] Joseph Lach. Letter: Baruch plan’s time has come. Bulletin of the Atomic Scientists, 46(5):42–43, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil90].

Lachow:2017:UDS

[Lac17] Irving Lachow. The upside and downside of swarming drones. Bulletin of the Atomic Scientists, 73(2):96–101, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1290879.

Ladendorf:1981:FRB

[Lad81] Robert Ladendorf. Film review: The Nuclear Watchdogs, produced by Esther Kartiganer. Bulletin of the Atomic Scientists, 37 (4):49–50, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFeber:1962:BBY

[LaF62] Walter LaFeber. Books: From Yalta to Disarmament; Cold War Debate, by Joseph Morray. Bulletin of the Atomic Scientists, 18(9):28, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFeber:1986:BRB

[LaF86] Walter LaFeber. Book review: Mayday: Eisenhower, Khrushchev and the U-2 Affair, by Michael R. Beschloss. Bulletin of the Atomic Scientists, 42(9):49–50, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFarge:1990:BRR

[LaF90a] Phyllis LaFarge. Book review: Rear-view strategy: Preventing World War III, by David Abshire. Bulletin of the Atomic Scientists, 46(5):38–39, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFollette:1990:LLW

[LaF90b] Melinda LaFollette. Letter: Lowered wall demands greater tolerance. Bulletin of the Atomic Scientists, 46(2):47, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaForge:1996:LIU

[LaF96] John M. LaForge. Letter: Invalid uses. Bulletin of the Atomic Scientists, 52(1):61, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nuc95].

Lago:1980:CGM

[Lag80] Don Lago. Commentary: A gust of madness. Bulletin of the Atomic Scientists, 36(9):50, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lago:1981:CLM

[Lag81] Don Lago. Commentary: On Lennon’s murder. Bulletin of the Atomic Scientists, 37(3):62–63, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lakoff:1964:TCP

[Lak64] Sanford A. Lakoff. The N th culture problem. Bulletin of the Atomic Scientists, 20(5):21–23, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lakoff:1966:BBR

[Lak66] Sanford A. Lakoff. Books: Tizard, by Ronald W. Clark. Forewords by Sir Solly Zuckerman and Vannevar Bush. Bulletin of the Atomic Scientists, 22(1):39–41, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lakoff:1978:CGR

[Lak78] Sanford A. Lakoff. Commentary: Grass roots TV looks at science and public policy. Bulletin of the Atomic Scientists, 34(10):51– 54, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [AL79a].

Lakoff:1982:CEP

[Lak82] Sanford A. Lakoff. Commentary: Energy: the price of development. Bulletin of the Atomic Scientists, 38(4):45–46, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:FND

[Lal64a] Arthur S. Lall. Forum: The nonaligned in disarmament negotiations. Bulletin of the Atomic Scientists, 20(5):17–21, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:DPP

[Lal64b] Betty Goetz Lall. Disarmament policy and the Pentagon. Bulletin of the Atomic Scientists, 20(7):37–40, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:FDI

[Lal64c] Betty Goetz Lall. Forum: On disarmament issues: Mainland China and U.S. security. Bulletin of the Atomic Scientists, 20(4): 24–27, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:IAC

[Lal64d] Betty Goetz Lall. Information in arms control verification. Bulletin of the Atomic Scientists, 20(8):43–45, October 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:NWD

[Lal64e] Betty Goetz Lall. A NATO–Warsaw detente? Bulletin of the Atomic Scientists, 20(9):37–39, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:NRC

[Lal64f] Betty Goetz Lall. News & reports: Congress considers impact of defense reductions. Bulletin of the Atomic Scientists, 20(2):31– 32, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:DID

[Lal64g] Betty Goetz Lall. On disarmament issues: Diminishing the danger of surprise attack in Europe. Bulletin of the Atomic Scientists, 20(3):31–34, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:DIPa

[Lal64h] Betty Goetz Lall. On disarmament issues: Peacekeeping since 1946. Bulletin of the Atomic Scientists, 20(5):37–40, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:DIPb

[Lal64i] Betty Goetz Lall. On disarmament issues: The Polish plan. Bulletin of the Atomic Scientists, 20(6):41–43, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1964:QAU

[Lal64j] Betty Goetz Lall. Questions and answers on the U.S. production freeze proposal. Bulletin of the Atomic Scientists, 20(10):30–34, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:EOP

[Lal65a] Arthur S. Lall. The explosion of October 16: The political effects of the Chinese bomb. Bulletin of the Atomic Scientists, 21(2):21– 24, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:AGR

[Lal65b] Betty Goetz Lall. Approaches to German reunification. Bulletin of the Atomic Scientists, 21(4):41–44, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:BBF

[Lal65c] Betty Goetz Lall. Books: The Future Character and Role of Peace Observation Arrangements Under the United Nations. Bulletin of the Atomic Scientists, 21(5):38–39, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:GSR

[Lal65d] Betty Goetz Lall. Government sponsored research for disarmament. Bulletin of the Atomic Scientists, 21(2):44–47, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:NAG

[Lal65e] Betty Goetz Lall. Notes from Asia and Germany. Bulletin of the Atomic Scientists, 21(9):33–35, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:PIA

[Lal65f] Betty Goetz Lall. Perspectives on inspection for arms control. Bulletin of the Atomic Scientists, 21(3):51–53, March 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1965:SRS

[Lal65g] Betty Goetz Lall. Substantial reductions in strategic delivery vehicles. Bulletin of the Atomic Scientists, 21(1):45–48, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:ACC

[Lal66a] Betty Goetz Lall. Arms control in Congress, 1966. Bulletin of the Atomic Scientists, 22(10):35–37, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:CDR

[Lal66b] Betty Goetz Lall. Conversion of defense resources. Bulletin of the Atomic Scientists, 22(1):46–48, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:CAC

[Lal66c] Betty Goetz Lall. Cooperation and arms control in outer space. Bulletin of the Atomic Scientists, 22(9):34–37, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:FNV

[Lal66d] Betty Goetz Lall. Forum: Nonintervention vs. containment. Bulletin of the Atomic Scientists, 22(5):21–24, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:ICW

[Lal66e] Betty Goetz Lall. International cooperation at the White House. Bulletin of the Atomic Scientists, 22(2):41–43, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:NFU

[Lal66f] Betty Goetz Lall. New frontiers of urban excellence? Bulletin of the Atomic Scientists, 22(3):37–40, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:RNP

[Lal66g] Betty Goetz Lall. Reports: Next: A proliferation ban. Bulletin of the Atomic Scientists, 22(1):42–43, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:RPU

[Lal66h] Betty Goetz Lall. Reports: Peacekeeping at the U.N. Bulletin of the Atomic Scientists, 22(8):43–45, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1966:RUC

[Lal66i] Betty Goetz Lall. Reports: U.S. China policy is changing: Travel comes first. Bulletin of the Atomic Scientists, 22(6):93–96, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:BBS

[Lal67a] Arthur S. Lall. Books: Neutralization in Southeast Asia, by Cyril E. Black, Richard A. Falk, Klaus Knorr, and Oran Young. Bulletin of the Atomic Scientists, 23(4):29–32, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:BBN

[Lal67b] Arthur S. Lall. Books: The Neutrals and the Test-Ban Negotiations, by M. Samir Ahmed. Bulletin of the Atomic Scientists, 23 (10):33–34, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:CDA

[Lal67c] Betty Goetz Lall. Congress debates the ABM. Bulletin of the Atomic Scientists, 23(7):28–32, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:GAD

[Lal67d] Betty Goetz Lall. Gaps in the ABM debate. Bulletin of the Atomic Scientists, 23(4):45–46, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:GCF

[Lal67e] Betty Goetz Lall. The Geneva Conference — five years later. Bulletin of the Atomic Scientists, 23(1):38–42, January 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:RAA

[Lal67f] Betty Goetz Lall. Reports: American attitudes on U.S.–Soviet relations. Bulletin of the Atomic Scientists, 23(1):34–38, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:SIU

[Lal67g] Betty Goetz Lall. Superiority and innovation in the U.S. defense forces. Bulletin of the Atomic Scientists, 23(3):11–13, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1967:UER

[Lal67h] Betty Goetz Lall. U.S.–European relations: Appraisal and future policy. Bulletin of the Atomic Scientists, 23(6):45–47, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1970:BBU

[Lal70] Betty Goetz Lall. Books: The United Nations: A View from Within, by Ralph Townley. Bulletin of the Atomic Scientists, 26(3):42–43, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lal:1975:LWO

[Lal75] Gobind Behari Lal. Letters: What is oriental? Bulletin of the Atomic Scientists, 31(6):2, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1977:MDN

[Lal77] Betty Goetz Lall. Mutual deterrence: the need for a new definition. Bulletin of the Atomic Scientists, 33(10):10–11, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lall:1986:DA

[Lal86] Betty G. Lall. Disarmament agency at 25. Bulletin of the Atomic Scientists, 42(10):30–32, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lamport:1969:BBM

[Lam69] James E. Lamport. Books: Appointment on the Moon, by Richard S. Lewis; The Promise of Space, by Arthur C. Clarke. Bulletin of the Atomic Scientists, 25(5):56, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaMarsh:1979:CWR

[LaM79] John R. LaMarsh. Commentary: What reactor failures? Bulletin of the Atomic Scientists, 35(4):55, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Atk78].

Lamarsh:1980:CNP

[Lam80] John R. Lamarsh. China’s nuclear power program. Bulletin of the Atomic Scientists, 36(5):28–31, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SL80].

Lamperti:1983:CRP

[Lam83] John Lamperti. Crisis relocation planning: “What harm can it do?”. Bulletin of the Atomic Scientists, 39(6):[various between 28–29], June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lampman:1985:LEN

[Lam85] Steve Lampman. Letter: European nonalignment. Bulletin of the Atomic Scientists, 41(10):55–56, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mea85, FUvH85].

Lamb:1987:CDR

[Lam87] John M. Lamb. Canadian defense: Roiling the arms control waters. Bulletin of the Atomic Scientists, 43(8):17–19, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1955:BRB

[Lan55] Alexander Langsdorf, Jr. Book review: The New Men, by C. P. Snow. Bulletin of the Atomic Scientists, 11(6):226, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1956:BBA

[Lan56a] Alexander Langsdorf, Jr. Books: Genetics in the Atomic Age, by Charlotte Auerbach. Bulletin of the Atomic Scientists, 12(9): 349, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1956:BBN

[Lan56b] Alexander Langsdorf, Jr. Books: The New Astronomy: The Physics and Chemistry of Life: First Book of Animals; Atomic Power; and Automatic Control, by the Editors of Scientific American. Bulletin of the Atomic Scientists, 12(3):93, March 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1957:BLN

[Lan57] Alexander Langsdorf, Jr. Books: The Laws of Nature, by R. E. Peierls. Bulletin of the Atomic Scientists, 13(2):72, February 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1958:BBL

[Lan58] Alexander Langsdorf, Jr. Books: Plant Life; The Universe; The New Chemistry; The Planet Earth; Lives in Science. Bulletin of the Atomic Scientists, 14(5):196, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Landsberg:1961:CMO

[Lan61a] H. E. Landsberg. Climate made to order. Bulletin of the Atomic Scientists, 17(9):370–374, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1961:TVB

[Lan61b] Alexander Langsdorf, Jr. Two views on Berlin. Bulletin of the Atomic Scientists, 17(7):294, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sti61].

Langsdorf:1962:AHC

[Lan62] Alexander Langsdorf, Jr. Arthur Holly Compton. Bulletin of the Atomic Scientists, 18(7):28–29, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Landis:1970:CSA

[Lan70] Elizabeth S. Landis. Communications: South African policy. Bulletin of the Atomic Scientists, 26(3):47, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langsdorf:1976:BBS

[Lan76] Alexander Langsdorf. Books: Beyond science and religion: The Universe of Experience, by Lancelot Law Whyte. Bulletin of the Atomic Scientists, 32(9):47, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1979:CWN

[Lan79] William J. Lanouette. Commentary from Washington: “No longer can the NRC say … ”. Bulletin of the Atomic Scientists, 35(6):6–8, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Col79a].

Lanouette:1980:KCR

[Lan80] William J. Lanouette. Kemeny Commission Report. Bulletin of the Atomic Scientists, 36(1):20–24, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1982:IPP

[Lan82a] William Lanouette. [Introduction to The peculiar politics of nuclear power]. Bulletin of the Atomic Scientists, 38(10):59, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1982:RNP

[Lan82b] William J. Lanouette. Reagan’s non-policy. Bulletin of the Atomic Scientists, 38(7):32–33, August/September 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1983:DM

[Lan83a] William Lanouette. Dream machine. Atlantic Monthly, ??(??):35– 52, 85–86, April 1983. ISSN 1072-7825 (print), 2151-9463 (electronic). A 20-page cover story about the history of the nuclear breeder reactor, a power plant designed to make more fuel than it consumes, including Szilard’s role both devising and naming the nuclear “chain reaction” in 1933 and the “breeder” in 1943. Letters and author’s replies in June, pages 6-7; July, pages 6 & 8; and August.

Lanouette:1983:CSS

[Lan83b] William J. Lanouette. Commentary: Safeguards secrecy. Bulletin of the Atomic Scientists, 39(6):51–52, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [dKL83].

Lancaster:1985:AEP

[Lan85] Carol J. Lancaster. Africa: Economics and politics of development. Bulletin of the Atomic Scientists, 41(8):27–30, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1989:BEN

[Lan89a] William Lanouette. Bulletin: Expectations for nonproliferation. Bulletin of the Atomic Scientists, 45(9):7, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1989:BHF

[Lan89b] William Lanouette. Bulletin: Hooked on foreign uranium. Bulletin of the Atomic Scientists, 45(8):4, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tyl89].

Lanouette:1989:BTB

[Lan89c] William Lanouette. Bumbling toward the bomb. Bulletin of the Atomic Scientists, 45(7):7–11, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). An article about Szilard, Einstein, and other refugee physicists alerting a suspicious U.S. Government about nuclear fission and nuclear weapons. Contains reproduction of August 2, 1939 Einstein/ Szilard letter to President Franklin D. Roosevelt.

Lanouette:1989:HCP

[Lan89d] William Lanouette. A history of cutoff proposals. Bulletin of the Atomic Scientists, 45(10):43, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano90-53].

Lanouette:1989:NTP

[Lan89e] William Lanouette. Nuclear testing: Public always the last to know. Bulletin of the Atomic Scientists, 45(8):11–12, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1989:PNS

[Lan89f] William Lanouette. Plutonium — no supply, no demand? Bulletin of the Atomic Scientists, 45(10):42–45, December 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1989:RRR

[Lan89g] William Lanouette. Report: Risk reduction: Toward better military manners. Bulletin of the Atomic Scientists, 45(10):8–9, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1989:TT

[Lan89h] William Lanouette. The tritium tiff. Bulletin of the Atomic Scientists, 45(10):44, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:ASWa

[Lan90a] William Lanouette. The Admiral and Star Wars. Bulletin of the Atomic Scientists, 46(1):38, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:ASWb

[Lan90b] William Lanouette. The Admiral and Stello wars. Bulletin of the Atomic Scientists, 46(1):40, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:BHE

[Lan90c] William Lanouette. Bulletin: Hands-on Energy [Department] leader. Bulletin of the Atomic Scientists, 46(4):3–4, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:BMM

[Lan90d] William Lanouette. Bulletin: Mickey Mouse, meet Andy Atom. Bulletin of the Atomic Scientists, 46(2):8, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi90c].

Lanouette:1990:BMC

[Lan90e] William Lanouette. Bulletin: Military or civilian superplane? Bulletin of the Atomic Scientists, 46(1):4, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:BNI

[Lan90f] William Lanouette. Bulletin: New, improved nukespeak. Bulletin of the Atomic Scientists, 46(1):5, January/February 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:BOS

[Lan90g] William Lanouette. Bulletin: One step forward, two steps back. Bulletin of the Atomic Scientists, 46(3):4, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:BPT

[Lan90h] William Lanouette. Bulletin: Painting themselves green. Bulletin of the Atomic Scientists, 46(5):4, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:GWH

[Lan90i] William Lanouette. Global warming: how much and why? Bulletin of the Atomic Scientists, 46(3):38–39, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:GSR

[Lan90j] William Lanouette. Greenhouse scare reheats nuclear debate. Bulletin of the Atomic Scientists, 46(3):34–37, April 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:JDW

[Lan90k] William Lanouette. James D. Watkins: Frustrated admiral of energy. Bulletin of the Atomic Scientists, 46(1):36–42, January/ February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:PEO

[Lan90l] William Lanouette. Photo essay: “Our town” v. “national security”. Bulletin of the Atomic Scientists, 46(10):30–36, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:RMB

[Lan90m] William Lanouette. Report: Military budget: The boom in B-2 bashing. Bulletin of the Atomic Scientists, 46(7):8–9, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:RSP

[Lan90n] William Lanouette. Report: Student Pugwash: Ideals for young scientists. Bulletin of the Atomic Scientists, 46(3):12–13, 52, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1990:WPS

[Lan90o] William Lanouette. Weapons plant at 40: Savannah River’s halo fades. Bulletin of the Atomic Scientists, 46(10):27–29, 37–38, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1992:ISP

[Lan92a] William Lanouette. Ideas by Szilard, physics by Fermi. Bulletin of the Atomic Scientists, 48(10):16–23, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special edition for the 50th Anniversary of the Chain Reaction (2 December 1942). An article about Leo Szilard’s role as co-designer of the world’s first nuclear reactor, and his invention of atomic secrecy.

Lanouette:1992:IS

[Lan92b] William Lanouette. Inventing secrecy. Bulletin of the Atomic Scientists, 48(10):19, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Landau:1996:JRR

[Lan96] Susan Landau. Joseph Rotblat: The road less traveled. Bulletin of the Atomic Scientists, 52(1):46–54, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1998:BRF

[Lan98] William Lanouette. Book review: Forgotten man: The Neutron and the Bomb: A Biography of Sir James Chadwick, by Andrew Brown. Bulletin of the Atomic Scientists, 54(1):61–62, January/ February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langeland:2001:HTE

[Lan01] Terje Langeland. Here, there, everywhere. Bulletin of the Atomic Scientists, 57(6):57–64, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/57.

Langeland:2002:MMP

[Lan02] Terje Langeland. Megatons to mega-problems. Bulletin of the Atomic Scientists, 58(3):49–56, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/3/49.

Lane:2008:SSP

[Lan08] Neal Lane. Science in the seat of power. Bulletin of the Atomic Scientists, 64(3):48, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/3/48.

Landau:2010:IQH

[Lan10] Emily B. Landau. The Iranian quagmire: How to move forward: Position: Tough bargaining tactics. Bulletin of the Atomic Scientists, 66(6):115–120, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/66/6/115.

Landau:2014:ZCW

[Lan14] Emily B. Landau. In the zone? Chemical weapons and the Middle East: The Israeli response. Bulletin of the Atomic Scientists, 70 (3):4–6, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lankov:2016:NKN

[Lan16] Andrei Lankov. North Korea, nuclear weapons, and the search for a new path forward. Bulletin of the Atomic Scientists, 72(5):340– ??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1948:ABE

[Lap48] Ralph E. Lapp. Atomic bomb explosions — effects on an American city. Bulletin of the Atomic Scientists, 4(2):49–54, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1950:SCD

[Lap50] Ralph E. Lapp. The strategy of civil defense. Bulletin of the Atomic Scientists, 6(8–9):241–243, August/September 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Lapp:1951:IDU

[Lap51a] Ralph E. Lapp. Industrial dispersion in the United States. Bulletin of the Atomic Scientists, 7(9):256–260, September 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1951:OMI

[Lap51b] Ralph E. Lapp. One missing item. Bulletin of the Atomic Scientists, 7(9):287, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1953:EEY

[Lap53] Ralph E. Lapp. Editorial: Eight years later. Bulletin of the Atomic Scientists, 9(7):234–236, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1954:IGV

[Lap54a] Ralph Lapp. An interview with Governor Val Peterson. Bulletin of the Atomic Scientists, 10(10):375–377, December 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1954:AC

[Lap54b] Ralph E. Lapp. Atomic candor. Bulletin of the Atomic Scientists, 10(8):312–314, 336, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1954:BBA

[Lap54c] Ralph E. Lapp. Books: The Atomic Submarine and Admiral Rickover, by Clay Blair, Jr. Bulletin of the Atomic Scientists, 10 (4):138, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1954:CDF

[Lap54d] Ralph E. Lapp. Civil defense faces new peril. Bulletin of the Atomic Scientists, 10(9):349–351, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1955:BRB

[Lap55a] Ralph E. Lapp. Book review: Atomic Weapons in Land Combat, by Col. G. C. Reinhardt and Lt. Col. W. R. Kintner. Bulletin of the Atomic Scientists, 11(3):93, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1955:FC

[Lap55b] Ralph E. Lapp. Fall-out and candor. Bulletin of the Atomic Scientists, 11(5):170, 200, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1955:GFI

[Lap55c] Ralph E. Lapp. Global fall-out [IV]. Bulletin of the Atomic Scientists, 11(9):339–343, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1955:RF

[Lap55d] Ralph E. Lapp. Radioactive fall-out. Bulletin of the Atomic Scientists, 11(2):45–51, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1955:RFI

[Lap55e] Ralph E. Lapp. Radioactive fall-out III. Bulletin of the Atomic Scientists, 11(6):206–209, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1955:RGL

[Lap55f] Ralph E. Lapp. Report from Geneva: The lesson of Geneva. Bulletin of the Atomic Scientists, 11(8):275, 308, October 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1956:BBE

[Lap56a] R. E. Lapp. Books: Nuclear Explosions and Their Effects. Bulletin of the Atomic Scientists, 12(10):377, December 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1956:HHB

[Lap56b] Ralph E. Lapp. The “Humanitarian” H-bomb. Bulletin of the Atomic Scientists, 12(7):261–264, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1956:LNP

[Lap56c] Ralph E. Lapp. A letter: Nuclear power secrecy. Bulletin of the Atomic Scientists, 12(4):135, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also [SW56].

Lapp:1956:SLP

[Lap56d] Ralph E. Lapp. Strontium limits in peace and war. Bulletin of the Atomic Scientists, 12(8):287–289, 320, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1957:BBE

[Lap57] Ralph E. Lapp. Books: Atomic Energy Technology for Lawyers, by E. Blythe Stason, Samuel D. Estep, and William J. Pierce; Workshops on Legal Problems of Atomic Energy. Bulletin of the Atomic Scientists, 13(5):179–180, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1958:CDS

[Lap58] Ralph E. Lapp. Civil defense shelters; an interview with Congressman Chet Holifield. Bulletin of the Atomic Scientists, 14(4): 130–134, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:BBD

[Lap59a] Ralph E. Lapp. Books: Basic Data Regarding Atomic Development Problems in Minnesota. Bulletin of the Atomic Scientists, 15(4):173, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:BBS

[Lap59b] Ralph E. Lapp. Books: Soviet Scientists Concerning the Dangers of Nuclear Weapons Tests. Bulletin of the Atomic Scientists, 15(8):349, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:CDS

[Lap59c] Ralph E. Lapp. Civil defense in the Soviet Union. Bulletin of the Atomic Scientists, 15(6):268, 270, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:FHD

[Lap59d] Ralph E. Lapp. Fallout and home defense. Bulletin of the Atomic Scientists, 15(5):187–191, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:FHS

[Lap59e] Ralph E. Lapp. Fallout hearings: Second round. Bulletin of the Atomic Scientists, 15(7):302–307, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:LFR

[Lap59f] Ralph E. Lapp. Locall fallout radioactivity. Bulletin of the Atomic Scientists, 15(5):181–186, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:RCC

[Lap59g] Ralph E. Lapp. Reader’s comment: A criticism of the GAC Report. Bulletin of the Atomic Scientists, 15(7):311–312, 320, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:SD

[Lap59h] Ralph E. Lapp. Sunshine and darkness. Bulletin of the Atomic Scientists, 15(1):27–29, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1959:WPN

[Lap59i] Ralph E. Lapp. What is the price of nuclear war? Bulletin of the Atomic Scientists, 15(8):340–343, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1960:LEG

[Lap60a] Ralph E. Lapp. Letter to the Editor: The GAC Report. Bulletin of the Atomic Scientists, 16(5):193, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wig60].

Lapp:1960:RCD

[Lap60b] Ralph E. Lapp. Rockefeller’s civil defense program. Bulletin of the Atomic Scientists, 16(4):134–136, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1961:REB

[Lap61a] R. E. Lapp. Review essay: Arms Control, Disarmament and National Security, by Donald G. Brennan. Bulletin of the Atomic Scientists, 17(7):287–289, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1961:NWS

[Lap61b] Ralph E. Lapp. Nuclear weapons systems. Bulletin of the Atomic Scientists, 17(3):99–102, 120, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1963:SO

[Lap63] Ralph E. Lapp. The strategy of overkill. Bulletin of the Atomic Scientists, 19(4):4–11, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1969:BBD

[Lap69] Ralph E. Lapp. Books: America Is In Danger, by General Curtis E. LeMay with Major General Dale O. Smith. Bulletin of the Atomic Scientists, 25(1):32–33, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1970:NWP

[Lap70] Ralph E. Lapp. Nuclear weapons: Past and present. Bulletin of the Atomic Scientists, 26(6):103–106, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1972:IAL

[Lap72] Ralph E. Lapp. Issues in arms limitations — four views: IV: SALT, MIRV and first-strike. Bulletin of the Atomic Scientists, 28(3):21–26, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1973:CC

[Lap73a] Ralph E. Lapp. The chemical century. Bulletin of the Atomic Scientists, 29(7):8–14, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapp:1973:VCS

[Lap73b] Ralph E. Lapp. A voice of conscience is stilled. III. Toward nuclear education. Bulletin of the Atomic Scientists, 29(6):6–8, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laponche:1976:RMF

[Lap76] Bernard Laponche. Rebuttal: Many in France oppose an ‘allnuclear’ policy. Bulletin of the Atomic Scientists, 32(10):44–45, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BPTB76a, BPTB76b].

Larsen:1950:GRC

[Lar50] Paul J. Larsen. The Government’s role in civil defense. Bulletin of the Atomic Scientists, 6(8–9):233–235, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Larkin:1962:GGH

[Lar62a] Timothy J. Larkin. The Great Ghat at Hardwar. Bulletin of the Atomic Scientists, 18(1):6–11, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Larrabee:1962:PS

[Lar62b] Eric Larrabee. The politics of strategy. Bulletin of the Atomic Scientists, 18(3):16–21, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Larson:1963:FMS

[Lar63] Arnold B. Larson. Forum: Modern science: The Old Adam. Bulletin of the Atomic Scientists, 19(7):28–29, September 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vL63].

Larus:1965:FTN

[Lar65] Joel Larus. Forum: Three on nuclear policies: III. To reduce the possibility of a nuclear catastrophe. Bulletin of the Atomic Scientists, 21(4):33–36, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lardner:1967:STA

[Lar67] Godfrey E. A. Lardner. Science and technology in Africa. Bulletin of the Atomic Scientists, 23(6):37–39, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dow67].

Larson:1984:CSR

[Lar84] Jeffry Larson. Commentary: Stockpile reliability. Bulletin of the Atomic Scientists, 40(7):56, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DM84].

Larson:1987:LDC

[Lar87] Jeffry Larson. Letter: The Daniloff case. Bulletin of the Atomic Scientists, 43(2):64, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha86].

Larrabee:1989:PSE

[Lar89] F. Stephen Larrabee. Perestroika shakes Eastern Europe. Bulletin of the Atomic Scientists, 45(2):25–29, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Larin:1999:MWW

[Lar99] Vladislav Larin. Mayak’s walking wounded. Bulletin of the Atomic Scientists, 55(5):20–27, September/October 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/5/20.

Larkin:2007:BFR

[Lar07] Emma Larkin. Burma’s forgotten refugees. Bulletin of the Atomic Scientists, 63(6):32–39, November/December 2007. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/63/6/32.

Lashof:1962:BBM

[Las62] Joyce C. Lashof. Books: May Man Prevail, by Erich Fromm. Bulletin of the Atomic Scientists, 18(9):28, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Latham:1973:LS

[Lat73] Gary V. Latham. Lunar seismology. Bulletin of the Atomic Scientists, 29(9):16–21, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lateef:1979:CPW

[Lat79] Noel V. Lateef. Commentary: Paving the way for technopolis. Bulletin of the Atomic Scientists, 35(9):58–61, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lattey:1984:CAI

[Lat84] Mike Lattey. Commentary: The arms industry. Bulletin of the Atomic Scientists, 40(3):64, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lattey:1991:LFD

[Lat91] Michael Lattey. Letter: Friends don’t give friends weapons. Bulletin of the Atomic Scientists, 47(6):45, July/August 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laurence:1946:BTP

[Lau46] William L. Laurence. The Bikini tests and public opinion. Bulletin of the Atomic Scientists, 2(5–6):2, 17, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laurence:1947:CPA

[Lau47] George C. Laurence. Canada’s participation in atomic energy development. Bulletin of the Atomic Scientists, 3(11):325–328, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laucks:1960:LET

[Lau60] Irving F. Laucks. Letter to the Editor: Too soon to move the clock? Bulletin of the Atomic Scientists, 16(5):193, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lauter:1964:LEC

[Lau64] Paul Lauter. Letter to the Editor: Courses for peace. Bulletin of the Atomic Scientists, 20(5):28, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laucks:1975:PWP

[Lau75] Irving F. Laucks. Peace work for the Pentagon. Bulletin of the Atomic Scientists, 31(8):37, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laucks:1976:LSF

[Lau76] Irving F. Laucks. Letter: Synthetic food. Bulletin of the Atomic Scientists, 32(7):5, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ben76a].

Laurin:1991:RD

[Lau91a] Fredrik Laurin. The report that disappeared. Bulletin of the Atomic Scientists, 47(2):14, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Laurin:1991:SUT

[Lau91b] Fredrik Laurin. Scandinavia’s underwater time bomb. Bulletin of the Atomic Scientists, 47(2):10–15, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zan91].

Lavrishchev:1955:SAP

[Lav55] A. N. Lavrishchev. Soviet atom pool. Bulletin of the Atomic Scientists, 11(8):279, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lawrence:1969:CCR

[Law69] John S. Lawrence. Current comments: Revolt against status quo. Bulletin of the Atomic Scientists, 25(3):31, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel68b, Fel69a].

Lawrence:1975:WHM

[Law75]  David Lawrence. What hath man wrought!  Bulletin of

the Atomic Scientists, 31(10):34–35, December 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from U.S. News & World Report, 17 August 1945.

Lawless:1984:CRW

[Law84a] W. F. Lawless. Commentary: Radioactive waste disposal. Bulletin of the Atomic Scientists, 40(7):55–56, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hun84].

Lawrence:1984:CDM

[Law84b] Mary B. Lawrence. Commentary: Dear members of the Board of Regents. Bulletin of the Atomic Scientists, 40(1):60–61, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author is the widow of Ernest O. Lawrence. She requests that the Board of Regents remove his name from Livermore weapons laboratory.

Lawless:1985:PMN

[Law85] William F. Lawless. Problems with military nuclear waste. Bulletin of the Atomic Scientists, 41(10):38–42, November 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lawless:1988:BRB

[Law88] William F. Lawless. Book review: Nuclear Imperatives and Public Trust: Dealing with Nuclear Waste, by Luther J. Carter. Bulletin of the Atomic Scientists, 44(8):48, 50, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lawrence:1990:LNE

[Law90] Regina G. Lawrence. Letter: Nicaraguan elections hardly “smooth”. Bulletin of the Atomic Scientists, 46(8):45, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kat90a].

Lawlor:2011:BSC

[Law11] Maj.Gen.Bruce Lawlor, (ret.). The Black Sea: Center of the nuclear black market. Bulletin of the Atomic Scientists, 67(6): 73–80, November/December 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/67/6/73.

Lay:1979:NTO

[Lay79] Fernando Lay. Nuclear technology in outer space. Bulletin of the Atomic Scientists, 35(7):27–31, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lay:2007:TSR

[Lay07] Jennie Lay. Test site rising. Bulletin of the Atomic Scientists, 63 (3):58–62, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/3/58.

Lapp:1953:CCS

[LB53] Ralph E. Lapp and Hanson Baldwin. Current comment: Sharing atomic developments with Britain; atomic secrecy. Bulletin of the Atomic Scientists, 9(3):91–92, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaFave:1968:BR

[LB68] Lawrence La Fave and J. G. Barense. Bertrand Russell. Bulletin of the Atomic Scientists, 24(6):56, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sto68a].

Libby:1978:FIU

[LB78a] Willard F. Libby and E. F. Black. Food irradiation: an unused weapon against hunger. Bulletin of the Atomic Scientists, 34(2): 50–55, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lipton-Barash:1978:LGS

[LB78b] Judith Lipton-Barash. Letter: Good suggestions. Bulletin of the Atomic Scientists, 34(8):63–64, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Law:2023:RWP

[LBCA23] Beverly Law, Ralph Bloemers, Nancy Colleton, and Mackenzie Allen. Redefining the wildfire problem and scaling solutions to meet the challenge. Bulletin of the Atomic Scientists, 79 (6):377–384, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2266941.

Lewis:1978:SRP

[LC78] John W. Lewis and Harlan Cleveland. Special report: Part 1: Peace and security in the Pacific: a Pacific settlement? Bulletin of the Atomic Scientists, 34(8):10–11, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Loeb:1988:LLF

[LC88] Benjamin S. Loeb and Malvine Cole. Letter: Leave “freeze” behind. Bulletin of the Atomic Scientists, 44(5):53–54, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fer88] and comment [Min88].

Lopez:1993:SDT

[LC93] George A. Lopez and David Cortright. Sanctions: Do they work? Bulletin of the Atomic Scientists, 49(9):14–15, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyman:1994:LGR

[LC94] Edwin S. Lyman and Luther J. Carter. Letters: Getting rid of weapon plutonium. Bulletin of the Atomic Scientists, 50(4): 59–60, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Car94a].

Lopez:1998:CD

[LC98a] George A. Lopez and David Cortright. Counting the dead. Bulletin of the Atomic Scientists, 54(3):41, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopez:1998:TGP

[LC98b] George A. Lopez and David Cortright. Trouble in the Gulf: Pain and promise. Bulletin of the Atomic Scientists, 54(3):39–43, May/ June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bos98, aI98].

Lindley:2009:LCN

[LC09] Dan Lindley and Kevin Clemency. Low-cost nuclear arms races. Bulletin of the Atomic Scientists, 65(2):44–51, March/April 2009.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/2/44.

Lund:1979:BRB

[LD79] Ebba Lund and Robert B. Dean. Book review: Water Re-Use and the Cities, by Roger E. Kasperson and Jeanne X. Kasperson. Bulletin of the Atomic Scientists, 35(7):46–48, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lapidus:1989:USR

[LD89] G. W. Lapidus and A. Dallin. U.S.–Soviet relations: The pacification of Ronald Reagan. Bulletin of the Atomic Scientists, 45(1):14–17, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyman:2000:TAP

[LD00] Edwin Lyman and Steven Dolley. The trouble: Accident prone. Bulletin of the Atomic Scientists, 56(2):42–46, March 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/2/42.

LeBert-Francis:1977:LRU

[Le 77] Brian Le Bert-Francis. Letter: Restructuring the U.N. Bulletin of the Atomic Scientists, 33(3):2–3, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hud76].

Leng:1985:CWR

[LE85] Russell J. Leng and William Epstein. Calculating weapons reductions. Bulletin of the Atomic Scientists, 41(2):39–41, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leaner:1947:PAW

[Lea47] Abba P. Leaner. The President addresses the world. Bulletin of the Atomic Scientists, 3(4–5):107–110, 126, April/May 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leavitt:1969:PAP

[Lea69] William Leavitt. Post-Apollo policy: A look into the 1970s. Bulletin of the Atomic Scientists, 25(7):41–44, September 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leaf:1970:SCN

[Lea70] Alexander Leaf. Social consequences of new developments in medicine. Bulletin of the Atomic Scientists, 26(1):21–22, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leaning:1987:SWR

[Lea87] Jennifer Leaning. Star Wars revives civil defense. Bulletin of the Atomic Scientists, 43(4):42–46, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leavitt:1989:PBC

[Lea89] Robert Leavitt. Perspective: Bush’s chintzy offer at Vienna. Bulletin of the Atomic Scientists, 45(8):13, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leahy:1997:GOD

[Lea97] Patrick Leahy. Guest opinion: December in Ottawa. Bulletin of the Atomic Scientists, 53(5):4, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leake:2007:MWF

[Lea07] Jonathan Leake. The man who flipped a nation. Bulletin of the Atomic Scientists, 63(2):24–29, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/2/24.

Lebow:1979:CTS

[Leb79] Richard Ned Lebow. Commentary: Thucydides’ speech to the American Senate. Bulletin of the Atomic Scientists, 35(10):7– 8, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lebow:1985:PWA

[Leb85] Richard Ned Lebow. Practical ways to avoid superpower crises. Bulletin of the Atomic Scientists, 41(1):22–28, January 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lebow:1987:DQL

[Leb87] Richard Ned Lebow. The dangers of quick launch. Bulletin of the Atomic Scientists, 43(9):36–39, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lebow:1988:RCC

[Leb88a] Richard Ned Lebow. A rejoinder: the case in not closed. Bulletin of the Atomic Scientists, 44(6):44, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gar88c].

Lebow:1988:WKB

[Leb88b] Richard Ned Lebow. Was Khrushchev bluffing in Cuba? Bulletin of the Atomic Scientists, 44(3):38–42, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gar88c, Leb88a].

Lebow:1991:BRB

[Leb91a] Richard Ned Lebow. Book review: Baiting Khrushchev: The Crisis Years, by Michael R. Beschloss. Bulletin of the Atomic Scientists, 47(10):41–42, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lebow:1991:BRK

[Leb91b] Richard Ned Lebow. Book reviews: Khrushchev redux: Chruschtchow: ein Politisches Portrait by Fedor Burlazki, Claasen, 1990; Khrushchev on Khrushchev: An Inside Account of the Man and His Era, by Sergei Khrushchev, translated by William Taubman, Lttle, Brown, 1990; Khrushchev Remembers: The Glasnost Tapes, by Nikita S. Khrushchev, Jerrold L. Schechter, and Vyacheslav V. Luchkov, editors, Little, Brown 1990. Bulletin of the Atomic Scientists, 47(4):43–45, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lebow:1992:BRC

[Leb92] Richard Ned Lebow. Book review: Camera on Cuba: Eyeball to Eyeball: The Inside Story of the Cuban Missile Crisis, by Dino A. Brugioni. Bulletin of the Atomic Scientists, 48(6):46, July/ August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lebow:1994:BRN

[Leb94] Richard Ned Lebow. Book review: Near-death experience: Cuba on the Brink: Castro, the Missile Crisis and the Soviet Collapse, edited by James G. Blight, Bruce J. Allyn, and David A. Welch. Bulletin of the Atomic Scientists, 50(4):55–56, July/ August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lederberg:1966:EGH

[Led66a] Joshua Lederberg. Experimental genetics and human evolution. Bulletin of the Atomic Scientists, 22(8):4–11, October 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Orn67, Led67].

Lederberg:1966:LEN

[Led66b] Joshua Lederberg. Letter to the Editor: A new proposal for European security. Bulletin of the Atomic Scientists, 22(10):28– 29, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lederberg:1967:FLR

[Led67] Joshua Lederberg. Forum: Lederberg replies. Bulletin of the Atomic Scientists, 23(6):60–61, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Led66a, Orn67].

Lederberg:1971:FMT

[Led71a] Joshua Lederberg. A freeze on missile testing. Bulletin of the Atomic Scientists, 27(3):4–6, 43, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lederberg:1971:SIC

[Led71b] Joshua Lederberg. Squaring an infinite circle: Radiobiology and the value of life. Bulletin of the Atomic Scientists, 27(7):43–45, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste72a] and reply [Led72].

Lederberg:1972:CR

[Led72] Joshua Lederberg. Communications: a reply. Bulletin of the Atomic Scientists, 28(6):5–6, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Led71b, Ste72a].

Lederman:1985:LSW

[Led85] Leon M. Lederman. Letter: Star Wars and Soviet technological lag. Bulletin of the Atomic Scientists, 41(10):54, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lederman:1994:PVP

[Led94] Leon Lederman. Paint a vision of the possible. Bulletin of the Atomic Scientists, 50(2):3, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lees:1960:LEI

[Lee60] Gerald M. Lees. Letter to the Editor: Ignorance knows least? Bulletin of the Atomic Scientists, 16(3):80, 113, March 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lee:1973:PES

[Lee73] Douglas H. K. Lee. Pollution and the environment: Specific approaches to health effects of pollutants. Bulletin of the Atomic Scientists, 29(8):45–47, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lee:1974:PUP

[Lee74] K. N. Lee. Public utilities and public policy: Water and politics in coastal California. Bulletin of the Atomic Scientists, 30(2):29– 35, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lee:1991:BBS

[Lee91a] Jennifer Scheck Lee. Bulletin: Red Star answers Bulletin. Bulletin of the Atomic Scientists, 47(3):5–6, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lee:1991:RRG

[Lee91b] Jennifer Scheck Lee. Renegade Russians grab for military control. Bulletin of the Atomic Scientists, 47(1):15–17, January 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lee:1997:SU

[Lee97] Rensselaer Lee. Smuggling update. Bulletin of the Atomic Scientists, 53(3):52–56, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lee:2024:DDP

[Lee24] Rachel Minyoung Lee. To do or not to do: Pyongyang’s seventh nuclear test calculations. Bulletin of the Atomic Scientists, 80(2):87–93, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2314434.

Lefkowitz:1976:LPR

[Lef76] Robert Lefkowitz. Letter: Pimping for the rich. Bulletin of the Atomic Scientists, 32(2):4, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pre75].

Lefkowitz:1977:LOD

[Lef77a] Louis J. Lefkowitz. A legal officer’s dilemma: the risks and implications of a scientific breakthrough. Bulletin of the Atomic Scientists, 33(5):11, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lefkowitz:1977:SNY

[Lef77b] Louis J. Lefkowitz. [Sidebar:] a New York State DNA Bill. Bulletin of the Atomic Scientists, 33(5):21, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leghorn:1956:PIS

[Leg56] Richard S. Leghorn. Part II: Science and military strategy: Controlling the nuclear threat in the second atomic decade. Bulletin of the Atomic Scientists, 12(6):189–195, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leghorn:1957:ARW

[Leg57] Richard S. Leghorn. The approach to a rational world security system. Bulletin of the Atomic Scientists, 13(6):195–200, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leghorn:1958:PAW

[Leg58] Richard S. Leghorn. The problem of accidental war. Bulletin of the Atomic Scientists, 14(6):205–209, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Legasov:1987:SED

[Leg87] Valery Legasov. A Soviet expert discusses Chernobyl. Bulletin of the Atomic Scientists, 43(6):32–34, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leggett:1992:BF

[Leg92a] Jeremy Leggett. Biological feedback. Bulletin of the Atomic Scientists, 48(5):31, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leggett:1992:GWW

[Leg92b] Jeremy Leggett. Global warming: the worst case. Bulletin of the Atomic Scientists, 48(5):28–33, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sch92a].

Leonard:1971:MPT

[LEH^{+}71] Eugene Leonard, Amitai Etzioni, Harvey A. Hornstein, Peter Abrams, Thomas Stephens, and Noel Tichy. MINERVA: A participatory technology system. Bulletin of the Atomic Scientists, 27(9):4–12, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ell72].

Leikind:1947:BAE

[Lei47] Morris C. Leikind. A bibliography of atomic energy (March 1, 1946–February 1, 1947). Bulletin of the Atomic Scientists, 3(4– 5):127–135, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leites:1968:FCR

[Lei68] Nathan Leites. Further comments on Reischauer and the choice on the war. Bulletin of the Atomic Scientists, 24(6):40–43, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leitenberg:1970:COS

[Lei70] Milton Leitenberg. Communications: Ocean science. Bulletin of the Atomic Scientists, 26(8):47–48, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leighton:1972:AVA

[Lei72a] Lauren G. Leighton. Another view of Akademgorodok. Bulletin of the Atomic Scientists, 28(4):37–42, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leitenberg:1972:PSW

[Lei72b] Milton Leitenberg. The present state of the world’s arms race. Bulletin of the Atomic Scientists, 28(1):15–21, January 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leich:1973:CMF

[Lei73] Harold H. Leich. Communications: Moon flights. Bulletin of the Atomic Scientists, 29(3):3, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Myr72].

Leitenberg:1974:ROS

[Lei74] Milton Leitenberg. The race to oblivion: The superpowers talk peace while preparing for war. Bulletin of the Atomic Scientists, 30(7):8–20, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano74c].

Leich:1975:SS

[Lei75] Harold H. Leich. Sewerless society. Bulletin of the Atomic Scientists, 31(9):38–44, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leitenberg:1978:LCI

[Lei78] Milton Leitenberg. Letter: The central issue of arms control and disarmament. Bulletin of the Atomic Scientists, 34(9):63– 64, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Whi79].

Leitenberg:1982:CSS

[Lei82a] Milton Leitenberg. The case of the stranded sub. Bulletin of the Atomic Scientists, 38(3):10–13, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [ML82] and erratum [Ano82g].

Leitenberg:1982:NGW

[Lei82b] Milton Leitenberg. The numbers game or ‘Who’s on first?’. Bulletin of the Atomic Scientists, 38(6):27–32, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BL83a].

Leitenberg:1983:CSV

[Lei83a] Milton Leitenberg. Commentary: On ‘seismic verification’. Bulletin of the Atomic Scientists, 39(6):52, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil83, Mar83d].

Leitenberg:1983:EDC

[Lei83b] Milton Leitenberg. From Europe: The dream of a common security. Bulletin of the Atomic Scientists, 39(9):34–36, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Eis84].

Leitenberg:1989:LSN

[Lei89a] Milton Leitenberg. Letter: Soviets never invaded anyone? Bulletin of the Atomic Scientists, 45(5):47, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gar88b].

Leitenberg:1989:LSS

[Lei89b] Milton Leitenberg. Letter: Sweden’s Soviet sub problem. Bulletin of the Atomic Scientists, 45(1):66, January/February 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano88c].

Leitenberg:1990:LUA

[Lei90] Milton Leitenberg. Letter: U.S. aid to Israel questioned. Bulletin of the Atomic Scientists, 46(4):52–53, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Sol90, Gel90].

Leitenberg:1992:BRD

[Lei92] Milton Leitenberg. Book review: Dangerous claims: Dangerous Liaison: The Inside Story of U.S.–Israeli Covert Relationship, by Andrew Cockburn and Leslie Cockburn. Bulletin of the Atomic Scientists, 48(2):42–43, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leiken:2007:GJ

[Lei07] Robert S. Leiken. Generation jihad. Bulletin of the Atomic Scientists, 63(3):63–65, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/3/63.

Lempert:1981:WYS

[Lem81] Robert Lempert. Will young scientists build bombs? Bulletin of the Atomic Scientists, 37(1):61–64, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lemarchand:1988:LBA

[Lem88] Guillermo A. Lemarchand. Letter: Buenos Aires oath. Bulletin of the Atomic Scientists, 44(6):54, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lentz:1960:LEF

[Len60] Theo. F. Lentz. Letter to the Editor: Further study needed. Bulletin of the Atomic Scientists, 16(1):0, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lenefsky:1973:NFU

[Len73] David Lenefsky. No-first-use of nuclear weapons — a pledge. Bulletin of the Atomic Scientists, 29(3):9, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lenske:1976:LNM

[Len76] Reuben Lenske. Letter: No moral right. Bulletin of the Atomic Scientists, 32(9):7, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Lens:1988:LMM

[Len88] Shirley Lens. Letter: Muste misperceived. Bulletin of the Atomic Scientists, 44(2):54, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sim87a].

Lentzos:2011:SBW

[Len11] Filippa Lentzos. Strengthening the biological weapons convention confidence-building measures: Toward a cycle of engagement. Bulletin of the Atomic Scientists, 67(3):26–33, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/3/26.

Lennane:2014:BBA

[Len14] Richard Lennane. Ban the bomb?: an Australian response. Bulletin of the Atomic Scientists, 70(6):10–13, November 2014. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lentzos:2020:HPW

[Len20] Filippa Lentzos. How to protect the world from ultra-targeted biological weapons. Bulletin of the Atomic Scientists, 76(6): 302–308, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1846412.

Lentzos:2021:MID

[Len21] Filippa Lentzos. Monitoring Iraq’s dual-use capabilities: an interview with Gabriele Kraatz-Wadsack. Bulletin of the Atomic Scientists, 77(4):172–176, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1941537.

Leontief:1964:FAA

[Leo64] Wassily Leontief. Forum: Alternatives to armament expenditures. Bulletin of the Atomic Scientists, 20(6):19–21, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leontief:1965:ITE

[Leo65] Wassily Leontief. An institute for technical economics. Bulletin of the Atomic Scientists, 21(7):46–47, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leong:2019:FRF

[Leo19] Brenda Leong. Facial recognition and the future of privacy: I always feel like … somebody’s watching me. Bulletin of the Atomic Scientists, 75(3):109–115, 2019. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1604886.

Lepper:1962:BBM

[Lep62] Mark H. Lepper. Books: The Medical Consequences of Thermonuclear War. Bulletin of the Atomic Scientists, 18(10):29–30, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lepgold:1989:BRB

[Lep89] Joseph Lepgold. Book review: The Other Side of Arms Control, by Alan B. Sherr. Bulletin of the Atomic Scientists, 45(4):46–47, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lerner:1946:HKP

[Ler46] A. P. Lerner. How to keep the peace … an analysis and a proposal. Bulletin of the Atomic Scientists, 2(5–6):3–4, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lerner:1948:EC

[Ler48] A. P. Lerner. An economist comments. Bulletin of the Atomic Scientists, 4(10):306–309, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a, Ash48, Rit48, Ano48v].

Lerner:1949:DCA

[Ler49a] A. P. Lerner. Does control of atomic energy involve a controlled economy? Bulletin of the Atomic Scientists, 5(1):15–16, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LeRoy:1949:AEL

[LeR49b] George V. LeRoy. “Atomic energy and the life sciences” AEC sixth semiannual report. Bulletin of the Atomic Scientists, 5(11): 315–319, 322, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LeRoy:1950:BRB

[LeR50a] George LeRoy. Book review: Atomic medicine, edited by C. F. Behrens. Bulletin of the Atomic Scientists, 6(5):160, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LeRoy:1950:CRH

[LeR50b] George V. LeRoy. Control of radiation hazards in the U.S. atomic energy program. Bulletin of the Atomic Scientists, 6(11):332, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leray:1952:SFE

[Ler52] Jean Leray. Some French experiences. Bulletin of the Atomic Scientists, 8(7):236–237, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Lerner:1953:SWB

[Ler53] Abba P. Lerner. Should we break our biggest monopoly? Bulletin of the Atomic Scientists, 9(4):110–114, 150–151, May 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [DP53].

Lerch:1980:BRB

[Ler80a] Irving A. Lerch. Book review: The Third World War August, 1985, by General Sir John Hackett and other NATO generals and advisors. Bulletin of the Atomic Scientists, 36(1):50–52, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lerch:1980:CUT

[Ler80b] Irving A. Lerch. Commentary: Unseen tremors. Bulletin of the Atomic Scientists, 36(8):62–63, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lerch:1981:CNP

[Ler81] Irving A. Lerch. Commentary: On the nuclear power debate: Fears and perceptions. Bulletin of the Atomic Scientists, 37 (9):49–52, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lessing:1963:SJC

[Les63] Lawrence Lessing. Science journalism: The coming age. Bulletin of the Atomic Scientists, 19(10):23, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lester:1981:SPN

[Les81] Richard Lester. Secrecy, patents and non-proliferation. Bulletin of the Atomic Scientists, 37(5):35–38, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leskov:1993:DGF

[Les93a] Sergei Leskov. Dividing the glory of the fathers. Bulletin of the Atomic Scientists, 49(4):37–39, May 1993. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

www.highbeam.com/doc/1G1-13869959.html.

Leskov:1993:LIS

[Les93b] Sergei Leskov. A lively interest in science history. Bulletin of the Atomic Scientists, 49(4):28, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hol94c].

Leskov:1993:NDS

[Les93c] Sergei Leskov. Notes from a dying spaceport. Bulletin of the Atomic Scientists, 49(8):40–43, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leskov:1993:RND

[Les93d] Sergei Leskov. Report: Nuclear dumping: Lies and incompetence. Bulletin of the Atomic Scientists, 49(5):13, 55, June 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leskov:1993:RUV

[Les93e] Sergei Leskov. Report: Ukraine: The view from Moscow. Bulletin of the Atomic Scientists, 49(9):8–10, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leskov:1994:SFU

[Les94] Sergei Leskov. The Sudoplatov file: An unreliable witness. Bulletin of the Atomic Scientists, 50(4):33–36, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction of date of Szilard picture [Ano94q].

Levi:1946:AEA

[Lev46] Edward H. Levi. The Atomic Energy Act: An analysis. Bulletin of the Atomic Scientists, 2(5–6):18–19, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levi:1947:SAE

[Lev47] Edward Levi. Shall the Atomic Energy Act be revised? Bulletin of the Atomic Scientists, 3(12):280, 285, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levi:1948:SAE

[Lev48] Edward Levi. Shall the Atomic Energy Act be revised? Bulletin of the Atomic Scientists, 4(9):280, 285, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1951:ISD

[Lev51a] Murray S. Levine. IV: Scientists have a duty to society. Bulletin of the Atomic Scientists, 7(12):376–378, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1951:NYC

[Lev51b] Murray S. Levine. The New York Committee on Atomic Information. Bulletin of the Atomic Scientists, 7(1):26–27, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1952:CDN

[Lev52] Murray S. Levine. Civil defense news. Bulletin of the Atomic Scientists, 8(2):62–63, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1953:CDV

[Lev53] Murray S. Levine. Civil defense vs. public apathy. Bulletin of the Atomic Scientists, 9(1):27–28, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levey:1960:LEO

[Lev60] J. S. Levey. Letter to the Editor: Oppenheimer and public disclosure. Bulletin of the Atomic Scientists, 16(4):145, April 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1962:BAS

[Lev62] Robert A. Levine. Breaking the arms stalemate. Bulletin of the Atomic Scientists, 18(6):8–11, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1963:UIC

[Lev63] Robert A. Levine. Unilateral initiatives: A cynic’s view. Bulletin of the Atomic Scientists, 19(1):22–25, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1964:OLM

[Lev64] Robert A. Levine. Open letter from a military intellectual to a sophisticated liberal leader. Bulletin of the Atomic Scientists, 20 (7):24–27, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BH65, Rav65, Lev65, Mos65].

Levine:1965:LR

[Lev65] Robert A. Levine. Levine replies. Bulletin of the Atomic Scientists, 21(3):33–34, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev64, Rav65].

Levin:1970:LIT

[Lev70] Nora Levin. Letter from Israel: Technical cooperation: Israel’s way in the Third World and Administered Territories. Bulletin of the Atomic Scientists, 26(4):46–52, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levin:1972:IDW

[Lev72a] Nora Levin. Israel and the developing world: New concepts in technical assistance. Bulletin of the Atomic Scientists, 28(9):37– 43, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1972:CEE

[Lev72b] Martin Levine. Communications: On the engineer in the establishment. Bulletin of the Atomic Scientists, 28(4):2, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Dru71, Dru72].

Levy:1973:CQI

[Lev73] Mark Levy. Communications: Queen Isabella and Mr. Stewart. Bulletin of the Atomic Scientists, 29(3):3, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste72b].

Levin:1976:LFF

[Lev76] Carl Levin. Letter: Far-fetched analogy. Bulletin of the Atomic Scientists, 32(9):5–6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Levato:1977:LFA

[Lev77a] Hugo Levato. Letter: Freedom in Argentina. Bulletin of the Atomic Scientists, 33(3):8, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lop76, Ama77].

Levick:1977:LSP

[Lev77b] Steve Levick. Letter: Suicidal patient. Bulletin of the Atomic Scientists, 33(10):2, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cla77].

Levick:1977:LWE

[Lev77c] Steve Levick. Letter: Worth exploring. Bulletin of the Atomic Scientists, 33(6):3, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [O’L77].

Levich:1978:A

[Lev78] Benjamin Levich. An appeal. Bulletin of the Atomic Scientists, 34(7):9, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levick:1980:CGS

[Lev80] Steve Levick. Commentary: Graphically speaking. Bulletin of the Atomic Scientists, 36(6):65, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leventhal:1984:WGS

[Lev84a] Paul Leventhal. From Washington: Getting serious about proliferation. Bulletin of the Atomic Scientists, 40(3):7–9, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levi:1984:BRB

[Lev84b] Barbara G. Levi. Book review: S.I.O.P.: The Secret US Plan for Nuclear War, by Peter Pringle and William Arkin. Bulletin of the Atomic Scientists, 40(9):50–52, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leventhal:1985:FNP

[Lev85] Paul Leventhal. Flaws in the Non-Proliferation Treaty. Bulletin of the Atomic Scientists, 41(8):12–15, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levin:1987:AWA

[Lev87] Carl Levin. Administration wrong on ABM Treaty. Bulletin of the Atomic Scientists, 43(3):30–33, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leventhal:1988:UJA

[Lev88] Paul Leventhal. U.S.–Japan accord invites proliferation. Bulletin of the Atomic Scientists, 44(4):11–13, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Levine:1989:ACC

[Lev89] Robert A. Levine. Arms control criteria: Informal moves work, too. Bulletin of the Atomic Scientists, 45(4):42–43, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leventhal:1990:CC

[Lev90] Paul Leventhal. Correction clarified. Bulletin of the Atomic Scientists, 46(4):53–54, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano90a].

Levin:1996:LTS

[Lev96] Ruth Levin. Letter: Try 30 seconds. Bulletin of the Atomic Scientists, 52(2):60–61, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leventhal:1998:RPM

[Lev98] Paul L. Leventhal. In response: A pox on MOX. Bulletin of the Atomic Scientists, 54(2):46–47, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98].

Levine:2006:IR

[Lev06] Haninah Levine. Improvised response. Bulletin of the Atomic Scientists, 62(4):22–23, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/4/22.

Levine:2007:BN

[Lev07] Haninah Levine. Birth of a notion. Bulletin of the Atomic Scientists, 63(4):38–40, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/4/38.

Levi:2012:SRV

[Lev12] Michael Levi. Splitting rock vs. splitting atoms: What shale gas means for nuclear power. Bulletin of the Atomic Scientists, 68 (4):52–60, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/4/52.

Lewis:1949:FRP

[Lew49] Layton Lewis. The fifth report: A press view. Bulletin of the Atomic Scientists, 5(3):93–94, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1950:CAE

[Lew50] W. Bennett Lewis. The Canadian Atomic Energy Project. Bulletin of the Atomic Scientists, 6(5):139–141, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1956:PIS

[Lew56] W. Arthur Lewis. Part IV: Science and world economy: Ten years of economic growth. Bulletin of the Atomic Scientists, 12 (8):290–295, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1961:LED

[Lew61a] David K. Lewis. Letter to the Editor: The danger of civil defense. Bulletin of the Atomic Scientists, 17(3):80, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1961:NNS

[Lew61b] W. Arthur Lewis. Needs of new states — science, men, and money. Bulletin of the Atomic Scientists, 17(2):43–47, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1962:BBN

[Lew62a] Richard Lewis. Books: Science and Nation, by J. Stefan Dupre and Sanford A. Lakoff. Bulletin of the Atomic Scientists, 18(5): 39, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1962:TFM

[Lew62b] Richard Lewis. Telstar: First with the most. Bulletin of the Atomic Scientists, 18(10):38–40, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1963:LPM

[Lew63a] Richard Lewis. The lessons of Project Mercury. Bulletin of the Atomic Scientists, 19(10):51–53, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1963:NRA

[Lew63b] Richard Lewis. News & report: Antarctic dilemma. Bulletin of the Atomic Scientists, 19(2):37–39, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1963:SS

[Lew63c] Richard Lewis. The SynCom Syndrome. Bulletin of the Atomic Scientists, 19(8):43–45, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1964:AC

[Lew64a] Richard Lewis. The Arizona compromise. Bulletin of the Atomic Scientists, 20(4):45–47, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1964:LQS

[Lew64b] Richard Lewis. … and quiet shines the sun. Bulletin of the Atomic Scientists, 20(1):30–32, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1964:SM

[Lew64c] Richard S. Lewis. At sea on the Moon. Bulletin of the Atomic Scientists, 20(9):35–37, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1965:MM

[Lew65a] Richard S. Lewis. The message from Mariner 4. Bulletin of the Atomic Scientists, 21(9):38–40, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1965:NWY

[Lew65b] Richard S. Lewis. Nawapa: Water for the year 2000. Bulletin of the Atomic Scientists, 21(5):9–11, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SSC68].

Lewis:1965:RMM

[Lew65c] Richard S. Lewis. Reports: The masterminds of Mars. Bulletin of the Atomic Scientists, 21(4):39–41, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1967:GNG

[Lew67a] Richard S. Lewis. Goal and no goal: A new policy in space. Bulletin of the Atomic Scientists, 23(5):17–20, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1967:PV

[Lew67b] Richard S. Lewis. A pennant on Venus. Bulletin of the Atomic Scientists, 23(9):19–24, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1967:SSP

[Lew67c] Richard S. Lewis. Science and space policy: Editorial introduction. Bulletin of the Atomic Scientists, 23(4):9–10, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1968:CSN

[Lew68a] Richard S. Lewis. Communiques: A streetcar named Apollo. Bulletin of the Atomic Scientists, 24(9):24, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1968:CSC

[Lew68b] Richard S. Lewis. Communiques: The Space Congress of Vienna. Bulletin of the Atomic Scientists, 24(9):25, November 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1968:CCF

[Lew68c] Richard S. Lewis. Current comments: The first men around the moon. Bulletin of the Atomic Scientists, 24(10):33, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1968:EA

[Lew68d] Richard S. Lewis. The end of Apollo. Bulletin of the Atomic Scientists, 24(7):4–6, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1968:KE

[Lew68e] Richard S. Lewis. The Kennedy effect. Bulletin of the Atomic Scientists, 24(3):2–5, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1968:PJ

[Lew68f] Richard S. Lewis. Panama junction. Bulletin of the Atomic Scientists, 24(5):12–16, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1969:ALL

[Lew69a] Richard S. Lewis. After the lunar landing: An introduction. Bulletin of the Atomic Scientists, 25(7):2, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1969:CCO

[Lew69b] Richard S. Lewis. Current comments: Our Terra–Luna transit system: Where will it take us? Bulletin of the Atomic Scientists, 25(3):22–23, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1969:SS

[Lew69c] Richard S. Lewis. The summer in space. Bulletin of the Atomic Scientists, 25(6):29–30, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1970:ARR

[Lew70a] Richard S. Lewis. Antarctic research and the relevance of science. Bulletin of the Atomic Scientists, 26(10):2–4, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Lewis:1970:ENL

[Lew70b] Richard S. Lewis. Evolution in NASA: Loss and cost of transition. Bulletin of the Atomic Scientists, 26(4):28–29, April 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1970:PPM

[Lew70c] Richard S. Lewis. A painless path to Mars. Bulletin of the Atomic Scientists, 26(1):44–45, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1970:SVW

[Lew70d] Richard S. Lewis. SALT in Vienna: The waltz of the powers. Bulletin of the Atomic Scientists, 26(7):19–21, September 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1970:WSC

[Lew70e] Richard S. Lewis. The Williamstown study of critical environmental problems. Bulletin of the Atomic Scientists, 26(8):24–30, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewontin:1970:JTT

[Lew70f] Richard C. Lewontin. The Jensen thesis: Three comments: 2. Further remarks on race and the genetics of intelligence. Bulletin of the Atomic Scientists, 26(5):23–25, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PTNR70].

Lewontin:1970:RI

[Lew70g] Richard C. Lewontin. Race and intelligence. Bulletin of the Atomic Scientists, 26(3):2–8, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1971:BSR

[Lew71a] Richard S. Lewis. A Bulletin special report: The radioactive salt mine. Bulletin of the Atomic Scientists, 27(6):27–30, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sku72].

Lewis:1971:CRS

[Lew71b] Richard S. Lewis. Comment: Requiem for the scientist– astronauts. Bulletin of the Atomic Scientists, 27(5):17–18, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1971:EAE

[Lew71c] Richard S. Lewis. End of Apollo: End of an era. Bulletin of the Atomic Scientists, 27(1):26–28, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1971:F

[Lew71d] Richard S. Lewis. Foreword. Bulletin of the Atomic Scientists, 27(7):2, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1971:RSA

[Lew71e] Richard S. Lewis. Reports: Student attitudes on science. Bulletin of the Atomic Scientists, 27(5):31–35, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1971:CCV

[Lew71f] W. Bennett Lewis. Communications: A convergent view. Bulletin of the Atomic Scientists, 27(5):48, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1972:BBN

[Lew72a] Richard S. Lewis. Books: Exploring New Ethics for Survival: The Voyage of the Spaceship Beagle, by Garrett Hardin. Bulletin of the Atomic Scientists, 28(9):45–46, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1972:ENP

[Lew72b] Richard S. Lewis. Editor’s note: [on publication schedule for One Dimensional Ecology]. Bulletin of the Atomic Scientists, 28(6):6, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1972:SSB

[Lew72c] Richard S. Lewis. From Seaborg to Schlesinger: A birdwatcher on the AEC. Bulletin of the Atomic Scientists, 28(4):44–45, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1972:RWB

[Lew72d] Richard S. Lewis. Report from Washington: Bailing out the breeder. Bulletin of the Atomic Scientists, 28(9):31–32, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1973:BBW

[Lew73a] Richard S. Lewis. Books: Stalking the Wild Taboo, by Garrett Hardin. Bulletin of the Atomic Scientists, 29(6):48, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1973:SDA

[Lew73b] Richard S. Lewis. The second discovery of America. Bulletin of the Atomic Scientists, 29(8):2, 28–31, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1976:LSD

[Lew76] Richard S. Lewis. Letter: Shock and dismay. Bulletin of the Atomic Scientists, 32(9):5, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Lewis:1980:CLWa

[Lew80a] Howard J. Lewis. Commentary: Letter from Washington. Bulletin of the Atomic Scientists, 36(4):13–16, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1980:CLWb

[Lew80b] Howard J. Lewis. Commentary: Letter from Washington. Bulletin of the Atomic Scientists, 36(5):15–18, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:2005:AA

[Lew05] Jeffrey Lewis. The ambiguous arsenal. Bulletin of the Atomic Scientists, 61(3):52–59, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/52.

Lewis:2007:CC

[Lew07] Jeffrey Lewis. A crisis of confidence. Bulletin of the Atomic Scientists, 63(1):13–15, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/13.

Lewis:2008:MD

[Lew08] Jeffrey Lewis. Minimum deterrence. Bulletin of the Atomic Scientists, 64(3):38–41, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/3/38.

Leys:1951:PPS

[Ley51] Wayne A. R. Leys. The political philosophy of scientists. Bulletin of the Atomic Scientists, 7(1):13–14, January 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [McR51].

Liepert:2015:GWA

[LG15] Beate G. Liepert and Alessandra Giannini. Global warming, the atmospheric brown cloud, and the changing Indian summer monsoon. Bulletin of the Atomic Scientists, 71(4):23–30, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lumpe:1992:TWM

[LGW92] Lora Lumpe, Lisbeth Gronlund, and David C. Wright. Third World missiles fall short. Bulletin of the Atomic Scientists, 48(2): 30–37, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Langway:1970:DTI

[LH70] C. C. Langway, Jr. and B. Lyle Hansen. Drilling through the ice cap: Probing climate for a thousand centuries. Bulletin of the Atomic Scientists, 26(10):62–66, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Leventhal:1988:LTF

[LH88] Paul Leventhal and Milton Hoenig. Letter: The tritium factor. Bulletin of the Atomic Scientists, 44(2):53–54, March 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AT88].

Leitenberg:1991:LSQ

[LH91] Milton Leitenberg and William D. Hartung. Letter: Small quibble over small arms. Bulletin of the Atomic Scientists, 47(3):47, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Har90d].

Luongo:1999:GIR

[LH99] Kenneth N. Luongo and William E. Hoehn III. Getting it right. Bulletin of the Atomic Scientists, 55(3):62–65, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/3/62.

Luongo:2005:OP

[LH05] Kenneth Luongo and William Hoehn. An ounce of prevention. Bulletin of the Atomic Scientists, 61(2):28–35, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/2/28.

Lewis:1970:CHI

[LI70] W. Bennett Lewis and David Rittenhouse Inglis. Communications: The hazardous industrial atom. Bulletin of the Atomic Scientists, 26(7):41, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ing70].

Libby:1948:RS

[Lib48] W. F. Libby. The radiocarbon story. Bulletin of the Atomic Scientists, 4(9):263–266, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Libby:1955:RF

[Lib55] Willard F. Libby. Radioactive fall-out. Bulletin of the Atomic Scientists, 11(7):256–260, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Libby:1957:LDL

[Lib57] Willard F. Libby. A letter from Dr. Libby to Dr. Schweitzer. Bulletin of the Atomic Scientists, 13(6):206–207, June 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch57b].

Libby:1958:RMN

[Lib58] W. F. Libby. Radiation and man: Nuclear tests and radiation: Distribution and effects of fall-out. Bulletin of the Atomic Scientists, 14(1):27–30, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Libby:1959:CT

[Lib59a] Willard F. Libby. The coming test. Bulletin of the Atomic Scientists, 15(4):172, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Libby:1959:NES

[Lib59b] Willard F. Libby. Nuclear energy — some new aspects. Bulletin of the Atomic Scientists, 15(6):240–244, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Libby:1965:MPP

[Lib65] Willard F. Libby. Man’s place in the physical Universe. Bulletin of the Atomic Scientists, 21(7):12–17, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Libicki:2001:DR

[Lib01] Martin C. Libicki. Defending the “Revolution”. Bulletin of the Atomic Scientists, 57(2):31–33, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/31.

Lichtenberg:1977:LSP

[Lic77] Don B. Lichtenberg. Letter: Soviet purity. Bulletin of the Atomic Scientists, 33(5):3, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Arb77].

Lie:1946:RUA

[Lie46] Trygve Lie. Report on the U.N. Atomic Energy Commission: Proceedings of the first meeting of the UN Atomic Energy Commission on June 13, 1946, and the second meeting of June 19, 1946. Bulletin of the Atomic Scientists, 2(1–2):2, 6–7, July 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lieberman:1962:LEE

[Lie62a] E. James Lieberman. Letter to the Editor: The ethical neutrality of LSD. Bulletin of the Atomic Scientists, 18(6):41, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LLH^{+}62].

Lieberman:1962:PW

[Lie62b] E. James Lieberman. Psychochemicals as weapons. Bulletin of the Atomic Scientists, 18(1):11–14, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lifton:1969:DLS

[Lif69] Robert Jay Lifton. Death in life: a statement. Bulletin of the Atomic Scientists, 25(6):39, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Comment on a award for the author’s book Death in Life: Survivors of Hiroshima.

Lifton:1976:NEW

[Lif76] Robert Jay Lifton. Nuclear energy and the wisdom of the body. Bulletin of the Atomic Scientists, 32(7):16–20, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [McN76].

Lifton:1980:PNW

[Lif80] Robert Jay Lifton. The prevention of nuclear war. Bulletin of the Atomic Scientists, 36(8):38–43, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BL81].

Lifton:1985:TNA

[Lif85] Robert Jay Lifton. Toward a nuclear-age ethos. Bulletin of the Atomic Scientists, 41(7):168–172, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lifton:2013:DCW

[Lif13] Robert Jay Lifton. The dimensions of contemporary war and violence: How to reclaim humanity from a continuing revolution in the technology of killing. Bulletin of the Atomic Scientists, 69(4): 9–17, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/69/
4/9.

Lifton:2023:OTO

[Lif23] Robert Jay Lifton. Oppenheimer’s tragedy — and ours. Bulletin of the Atomic Scientists, 79(4):216–220, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2223073.

Lightman:1984:PLS

[Lig84] Alan P. Lightman. Perspective: Lost in space. Bulletin of the Atomic Scientists, 40(5):2–3, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gre84b].

Lilienthal:1946:HCA

[Lil46] David E. Lilienthal. How can atomic energy be controlled? Bulletin of the Atomic Scientists, 2(7–8):14–15, 18, October 1, 1946.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1947:AEA

[Lil47a] David E. Lilienthal. Atomic energy and American industry. Bulletin of the Atomic Scientists, 3(11):339–340, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1947:AEY

[Lil47b] David E. Lilienthal. Atomic energy is Your business. Bulletin of the Atomic Scientists, 3(11):335–338, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1947:OAI

[Lil47c] David E. Lilienthal. Organization and administration of the International Agency. Bulletin of the Atomic Scientists, 3(8):203– 206, 231–232, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1948:PEP

[Lil48] David E. Lilienthal. Public employment or public pillory? Bulletin of the Atomic Scientists, 4(10):293–294, October 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1949:ARI

[Lil49a] David E. Lilienthal. AEC reply to industrial advisors. Bulletin of the Atomic Scientists, 5(2):56, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AEC49].

Lilienthal:1949:PIP

[Lil49b] David E. Lilienthal. Private industry and the public atom. Bulletin of the Atomic Scientists, 5(1):6–8, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1949:SSM

[Lil49c] David E. Lilienthal. Science and the spirit of man. Bulletin of the Atomic Scientists, 5(4):98–100, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1949:WDW

[Lil49d] David E. Lilienthal. Where do we go from here? Bulletin of the Atomic Scientists, 5(11):294, 308, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1953:CCS

[Lil53] David E. Lilienthal. Current comment: Secrecy or candor. Bulletin of the Atomic Scientists, 9(10):378–379, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpted from the author’s article The Case for Candor on National Security, published in the New York Times, 4 October 1953.

Lilienthal:1962:CRE

[Lil62] David E. Lilienthal. [Controlled release of energy]. Bulletin of the Atomic Scientists, 18(10):16, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lilienthal:1963:FMA

[Lil63] David E. Lilienthal. Forum: The mythology of the atom. Bulletin of the Atomic Scientists, 19(8):21–24, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha64, Bar64b, Pis64].

Line:1955:LEE

[Lin55] Paul B. Line. Letter to the Editor: Evacuation from our cities. Bulletin of the Atomic Scientists, 11(5):201, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindbeck:1969:ISP

[Lin69] John M. H. Lindbeck. An isolationist science policy. Bulletin of the Atomic Scientists, 25(2):66–72, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Linowitz:1973:FA

[Lin73] Sol M. Linowitz. The future of the Americas. Bulletin of the Atomic Scientists, 29(8):4–6, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Linebaugh:1979:CES

[Lin79] David Linebaugh. Commentary: European security: three proposals. Bulletin of the Atomic Scientists, 35(5):11–14, May 1979.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lin:1982:CDM

[Lin82] Herbert Lin. Commentary: Deterrence: MAD or NUT. Bulletin of the Atomic Scientists, 38(4):54–55, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindell:1984:BRB

[Lin84a] Erik Lindell. Book review: NATO’s Theater Nuclear Force Modernization Program: The Real Issue, by Jeffrey Record. Bulletin of the Atomic Scientists, 40(3):46–47, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Linebaugh:1984:WIS

[Lin84b] Dave Linebaugh. From Washington: INF + START = negotiations. Bulletin of the Atomic Scientists, 40(4):11–12, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Linebaugh:1984:PAC

[Lin84c] David Linebaugh. Perspective: Arms control priorities for the next Administration. Bulletin of the Atomic Scientists, 40(7):3– 4, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Linenthal:1984:RS

[Lin84d] Edward Tabor Linenthal. Religious studies. Bulletin of the Atomic Scientists, 40(10):15s–16s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindop:1985:BRB

[Lin85] Patricia J. Lindop. Book review: Nuclear Radiation: Risks and Benefits, by Edward Pochin. Bulletin of the Atomic Scientists, 41(6):48–49, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lin:1986:NTC

[Lin86] Herbert Lin. “No technology can support Reagan’s goal”. Bulletin of the Atomic Scientists, 42(2):35, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindgren:1988:CSO

[Lin88] David T. Lindgren. Commercial satellites open skies. Bulletin of the Atomic Scientists, 44(3):34–37, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Str88].

Lindsay:2008:GS

[Lin08] Stuart Lindsay. Genetic sequencing. Bulletin of the Atomic Scientists, 64(2):50–53, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/64/2/50.

Lin:2012:VNS

[Lin12] Herbert Lin. A virtual necessity: Some modest steps toward greater cybersecurity. Bulletin of the Atomic Scientists, 68(5): 75–87, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/68/5/75.

Lin:2019:ETC

[Lin19] Herbert Lin. The existential threat from cyber-enabled information warfare. Bulletin of the Atomic Scientists, 75(4):187–196, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1629574.

Lippman:1954:OC

[Lip54] Walter Lippman. The Oppenheimer case. Bulletin of the Atomic Scientists, 10(5):183, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lipson:1961:GPF

[Lip61] Leon Lipson. The Gagarin and Powers flights. Bulletin of the Atomic Scientists, 17(7):274, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lisco:1946:RHR

[Lis46a] Hermann Lisco. Radiation hazards and radiation sickness. Bulletin of the Atomic Scientists, 2(9–10):26–27, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lister:1946:GPR

[Lis46b] Gordon K. Lister. Government patent rights. Bulletin of the Atomic Scientists, 1(11):15–16, May 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lisco:1958:BBW

[Lis58] Hermann Lisco. Books: Radiation: What It Is and How It Affects You, by Jack Schubert and Ralph E. Lapp. Bulletin of the Atomic Scientists, 14(2):93, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Littlejohn:1954:BBH

[Lit54] C. Littlejohn. Books: A History of the Theories of Aether and Electricity (1900–1926), by Sir Edmund Whittaker. Bulletin of the Atomic Scientists, 10(6):233, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Little:1978:LUR

[Lit78] Winston W. Little. Letter: Uninspected reactors. Bulletin of the Atomic Scientists, 34(4):3, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dah78].

Livingston:1951:BOC

[Liv51] Goodhue Livingston, Jr. The blight of our cities. Bulletin of the Atomic Scientists, 7(9):260–262, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Livingston:1956:IS

[Liv56] M. Stanley Livingston. Information security. Bulletin of the Atomic Scientists, 12(5):172–174, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Livingston:1963:PC

[Liv63] Robert Livingston. Perception and commitment. Bulletin of the Atomic Scientists, 19(2):14–18, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Livingston:1968:ILS

[Liv68] Dennis Livingston. An international law of science: Orders on man’s expanding frontiers. Bulletin of the Atomic Scientists, 24 (10):6–10, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Livingston:1969:BBB

[Liv69] M. Stanley Livingston. Books: The Big Machine, by Robert Jungk. Bulletin of the Atomic Scientists, 25(3):34–35, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Luik:2017:PSN

[LJ17] Jüri Luik and Tomas Jermalavicius. A plausible scenario of nuclear war in Europe, and how to deter it: A perspective from Estonia. Bulletin of the Atomic Scientists, 73(4):233–239, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1338014.

Lorenz:2010:UFI

[LK10] Thomas Lorenz and Joanna Kidd. An uncertain future for international fuel banks. Bulletin of the Atomic Scientists, 66(3):44– 49, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
3/44.

Lightfoot:1989:LPT

[LKWR89] Frederick S. Lightfoot, Herbert Kriedman, Lawrence Wichlinski, and Jay Rosen. Letters: “Phantoms” talk back. Bulletin of the Atomic Scientists, 45(7):45–46, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ros89a].

Leaning:1983:PSN

[LL83] Jennifer Leaning and Matthew Leighton. Programs for surviving nuclear war: a critique. Bulletin of the Atomic Scientists, 39 (6):[various between 28–29], June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lach:1990:BRC

[LL90] Joseph Lach and Michael Lach. Book review: The conscience of science: The privilege of being a physicist, by Victor F. Weisskopf. Bulletin of the Atomic Scientists, 46(7):42–43, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyman:1997:BS

[LL97] Edwin S. Lyman and Paul Leventhal. Bury the stuff. Bulletin of the Atomic Scientists, 53(2):45–48, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See clarification [Lym97].

Lewis:2012:MCN

[LL12] John W. Lewis and Xue Litai. Making China’s nuclear war plan. Bulletin of the Atomic Scientists, 68(5):45–65, September/ October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/68/
5/45.

Lewis:2016:CSA

[LL16] John W. Lewis and Xue Litai. China’s security agenda transcends the South China Sea. Bulletin of the Atomic Scientists, 72(4): 212–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Llano:1970:SAB

[Lla70] George A. Llano. A survey of Antarctic biology: Life below freezing. Bulletin of the Atomic Scientists, 26(10):67–74, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Leary:1962:PNS

[LLH^{+}62] Timothy Leary, George Litwin, Michael Hollingshead, Gunther Weil, and Richard Alpert. The politics of the nervous system. Bulletin of the Atomic Scientists, 18(5):26–27, May 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [Lie62a, FL62].

Lloyd:1971:SAH

[Llo71] William Bross Lloyd, Jr. Switzerland: Absent host to the United Nations. Bulletin of the Atomic Scientists, 27(1):42–43, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1947:HLD

[LLR^{+}47] Lord Archbishop of York, F. A. Lindemann, Bertrand Russell, Lord Lindsay of Birker, and Lord Strabolgi. The House of Lords debates the control of atomic energy. Bulletin of the Atomic Scientists, 3(7):183–186, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovins:1983:CCT

[LLS83] Amory B. Lovins, L. Hunter Lovins, and Bernard L. Spinrad. Commentary: ‘The connection is tenuous’. Bulletin of the Atomic Scientists, 39(5):62–64, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spi83].

Lluma:1999:TLP

[Llu99] Diego Llumá. Terrorism: Low probability, high consequence. Bulletin of the Atomic Scientists, 55(6):14–16, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/6/14.

Lluma:2000:FSU

[Llu00a] Diego Llumá. Former Soviet Union: What the Russians left behind. Bulletin of the Atomic Scientists, 56(3):14–17, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/3/14.

Lluma:2000:WRL

[Llu00b] Diego Lluma. What the Russians left behind. Bulletin of the Atomic Scientists, 56(3):14, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyons:1961:LER

[LM61a] Gene M. Lyons and Louis Morton. Letter to the Editor: Reply. Bulletin of the Atomic Scientists, 17(5–6):255–256, May/ June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LM61b, Hah61].

Lyons:1961:SS

[LM61b] Gene M. Lyons and Louis Morton. School for strategy. Bulletin of the Atomic Scientists, 17(3):103–106, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Hah61] and reply [LM61a]. See also [SH61, Mor61b].

Lewis:1971:SCV

[LM71] W. Bennett Lewis and A. M. Marko. Seeking a convergent view: The unfinished detective story. Bulletin of the Atomic Scientists, 27(9):50–52, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Leitenberg:1975:LOD

[LM75] Milton Leitenberg and Matthew Melko. Letters: Obscene definition. Bulletin of the Atomic Scientists, 31(7):2–3, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mel75].

Lewis:1976:EMT

[LM76] John M. Lewis and Douglas P. Murray. Editorial: Mao Tse-Tung (1893–1976). Bulletin of the Atomic Scientists, 32(9):2, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lamarsh:1980:WPF

[LM80] John R. Lamarsh and Marvin M. Miller. Weapons proliferation and foreign students. Bulletin of the Atomic Scientists, 36(3): 25–30, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Linebaugh:1983:WNB

[LN83] David Linebaugh and Joseph Pepe Newman. From Washington: No bargains in Geneva. Bulletin of the Atomic Scientists, 39(7): 9–10, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindee:1986:CHC

[LN86] Susan Lindee and Dorothy Nelkin. Challenger: the high cost of hype. Bulletin of the Atomic Scientists, 42(9):16–18, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1989:BGK

[Loc89a] Dunbar Lockwood. Bulletin: Grudging kindness, gentle exaggeration. Bulletin of the Atomic Scientists, 45(10):4–5, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1989:LCW

[Loc89b] Dunbar Lockwood. Letter: Cruise was once dead. Bulletin of the Atomic Scientists, 45(6):51, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa89m].

Lockwood:1991:RSW

[Loc91a] Dunbar Lockwood. Report: Soviet weapons: ‘Modernization’ means cuts and slowdowns. Bulletin of the Atomic Scientists, 47(5):8–9, 47, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1991:RSM

[Loc91b] Dunbar Lockwood. Report: START: Missile-aneous issues settled at last. Bulletin of the Atomic Scientists, 47(7):12–13, 46, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1992:RAS

[Loc92a] Dunbar Lockwood. Report: After START: Bidding down. Bulletin of the Atomic Scientists, 48(3):8–10, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1992:RSI

[Loc92b] Dunbar Lockwood. Report: START II: The pen-chant for peace. Bulletin of the Atomic Scientists, 48(8):10–11, 45, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1993:DAR

[Loc93a] Dunbar Lockwood. Dribbling aid to Russia. Bulletin of the Atomic Scientists, 49(6):39–42, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1993:PNL

[Loc93b] Dunbar Lockwood. Parceling out Nunn–Lugar. Bulletin of the Atomic Scientists, 49(6):40–41, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1993:RAC

[Loc93c] Dunbar Lockwood. Report: Arms control: On Clinton’s calendar. Bulletin of the Atomic Scientists, 49(1):6–8, January/ February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1994:RDP

[Loc94] Dunbar Lockwood. Report: Disarmament: Purchasing power. Bulletin of the Atomic Scientists, 50(2):10–12, March/April 1994.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lockwood:1995:RNL

[Loc95] Dunbar Lockwood. Report: Nunn–Lugar: Getting down to business. Bulletin of the Atomic Scientists, 51(1):12–13, January/ February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lochbaum:2014:LAN

[Loc14] David Lochbaum. Life after nuclear: Decommissioning power reactors. Bulletin of the Atomic Scientists, 70(4):26–36, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lodgaard:1984:BND

[Lod84] Sverre Lodgaard. Breaking the nuclear deadlock: Nuclear disengagement in Europe. Bulletin of the Atomic Scientists, 40(3): 53–55, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Loeb:1984:ACD

[Loe84] Benjamin S. Loeb. On The Arms Control and Disarmament Agency. Bulletin of the Atomic Scientists, 40(6):54–55, June/ July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi84b].

Loeb:1985:LNB

[Loe85] Benjamin S. Loeb. Letter: Nagasaki bombing. Bulletin of the Atomic Scientists, 41(9):56, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Loeb:1986:LSN

[Loe86] Benjamin S. Loeb. Letter: Superpowers’ NPT bargain. Bulletin of the Atomic Scientists, 42(4):55, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nat86b].

Loeb:1987:ACT

[Loe87] Benjamin S. Loeb. Amend the Constitution’s treaty clause. Bulletin of the Atomic Scientists, 43(8):38–41, October 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho87].

Loeb:1988:LNN

[Loe88] Benjamin S. Loeb. Letter: The nine numbers. Bulletin of the Atomic Scientists, 44(3):57, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HO88].

Loeb:1989:LNF

[Loe89] Benjamin S. Loeb. Letter: No fast treaties. Bulletin of the Atomic Scientists, 45(3):44–45, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gad89].

Loeb:1995:LWN

[Loe95] Benjamin S. Loeb. Letter: But why Nagasaki? Bulletin of the Atomic Scientists, 51(1):3, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mey94].

Loescher:2002:BVR

[Loe02] Gil Loescher. Blaming the victim: Refugees and global security. Bulletin of the Atomic Scientists, 58(6):46–53, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/6/46.

Lofton:1947:ASA

[Lof47] John Lofton. Association of Scientists for Atomic Education. Bulletin of the Atomic Scientists, 3(7):196, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Loftus:1951:BRE

[Lof51] Joseph E. Loftus. Book review: Economic Aspects of Atomic Power, by S. H Schurr and J. Marschak. Bulletin of the Atomic Scientists, 7(3):70–74, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lofton:1960:SNS

[Lof60] John Lofton. A scientist’s non-scientific observations. Bulletin of the Atomic Scientists, 16(5):186–187, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ore60a].

Loftin:1984:CIR

[Lof84] Robert W. Loftin. Commentary: Ionizing radiation. Bulletin of the Atomic Scientists, 40(3):58–59, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [HQL84].

Longarzo:1950:RWC

[Lon50] L. C. Longarzo. Religious and Welfare Committee of NYCAI. Bulletin of the Atomic Scientists, 6(2):60, February 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lonsdale:1951:EPS

[Lon51] Kathleen Lonsdale. The ethical problems of science. Bulletin of the Atomic Scientists, 7(7–8):201–204, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ma’52].

London:1952:SPP

[Lon52a] Ivan D. London. Soviet psychology and psychiatry. Bulletin of the Atomic Scientists, 8(3):70–73, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano52f].

Longair:1952:LEU

[Lon52b] A. K. Longair. Letter to the Editor: Utilization of Harwell reactor heat to warm buildings. Bulletin of the Atomic Scientists, 8(2): 65, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lonsdale:1952:SPI

[Lon52c] Kathleen Lonsdale. A symposium: Part II: The state of Russian science today. III: Impressions from a recent Moscow visit. Bulletin of the Atomic Scientists, 8(3):66–70, 78, March 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

London:1957:TRA

[Lon57] Ivan D. London. Toward a realistic appraisal of Soviet science. Bulletin of the Atomic Scientists, 13(5):169–173, 176, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lonsdale:1958:SP

[Lon58] Kathleen Lonsdale. Scientists and the people. Bulletin of the Atomic Scientists, 14(7):242–245, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1964:IST

[Lon64] Franklin A. Long. The immediate steps toward disarmament. Bulletin of the Atomic Scientists, 20(2):7–10, February 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1967:SFA

[Lon67] F. A. Long. Scientists in foreign affairs: Where do we go now? Bulletin of the Atomic Scientists, 23(3):14–18, March 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1968:SBA

[Lon68] Franklin A. Long. Strategic balance and the ABM. Bulletin of the Atomic Scientists, 24(10):2–5, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1969:IIA

[Lon69a] Franklin A. Long. The industrial impact of Apollo. Bulletin of the Atomic Scientists, 25(7):70–73, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lonsdale:1969:DNS

[Lon69b] Dame Kathleen Lonsdale. Developing nations and scientific responsibility. Bulletin of the Atomic Scientists, 25(9):27–28, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1973:PNR

[Lon73] F. A. Long. President Nixon’s 1973 reorganization plan no. 1: Where do science and technology go now? Bulletin of the Atomic Scientists, 29(5):5–8, 40–42, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lonsdale:1974:CTI

[Lon74] H. K. Lonsdale. Communications: Towing icebergs. Bulletin of the Atomic Scientists, 30(3):2, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WC73, WC74].

Long:1975:ESW

[Lon75] F. A. Long. Editorial: Should we buy the Vladivostok agreement? Bulletin of the Atomic Scientists, 31(2):5–6, February 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1976:PNE

[Lon76a] F. A. Long. Peaceful nuclear explosions. Bulletin of the Atomic Scientists, 32(8):18–28, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [You77a].

Long:1976:PBO

[Lon76b] F. A. Long. The product has been oversold. Bulletin of the Atomic Scientists, 32(8):27, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1978:BBE

[Lon78] T. Dixon Long. Books: The Electric War: The Fight Over Nuclear Power, by Sheldon Novick. Bulletin of the Atomic Scientists, 34(4):59–60, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1982:BRB

[Lon82] Franklin A. Long. Book review: Kennedy, Khrushchev, and the Test Ban, by Glenn T. Seaborg with the assistance of Benjamin S. Loeb. Bulletin of the Atomic Scientists, 38(6):47–48, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1984:BND

[Lon84] Franklin A. Long. Breaking the nuclear deadlock: Unilateral initiatives. Bulletin of the Atomic Scientists, 40(5):50–54, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Long:1985:LMN

[Lon85a] John E. Long. Letter: Managing nuclear forces. Bulletin of the Atomic Scientists, 41(6):55, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil85a].

Longstreth:1985:RAS

[Lon85b] Thomas K. Longstreth. Report aims to sabotage arms control. Bulletin of the Atomic Scientists, 41(1):29–32, January 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [You85a].

Longworth:1985:BRB

[Lon85c] Richard Longworth. Book review: Nomenklatura: The Soviet Ruling Class, An Insider’s Report, by Michael Voslensky. Bulletin of the Atomic Scientists, 41(4):56–58, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wei86c].

Long:1986:GDU

[Lon86a] Franklin A. Long. Government dollars for university research. Bulletin of the Atomic Scientists, 42(3):45–49, March 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longworth:1986:RMT

[Lon86b] Richard L. Longworth. Reporting from Moscow: then and now. Bulletin of the Atomic Scientists, 42(10):20–21, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longstreth:1987:WRL

[Lon87] Thomas K. Longstreth. Washington report: Latest ABM ploy — old is new. Bulletin of the Atomic Scientists, 43(10):3–4, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longworth:1992:EEP

[Lon92] Richard C. Longworth. Eastern Europe: The party’s over. Bulletin of the Atomic Scientists, 48(1):22–29, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longworth:1995:PFD

[Lon95] Richard C. Longworth. Phantom forces, diminished dreams. Bulletin of the Atomic Scientists, 51(2):24–28, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longworth:1998:BB

[Lon98] Richard C. Longworth. Boomtown Baku. Bulletin of the Atomic Scientists, 54(3):34–38, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longworth:1999:NKW

[Lon99] R. C. Longworth. A new kind of war. Bulletin of the Atomic Scientists, 55(4):28–31, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/4/28.

Loomis:1950:CPS

[Loo50] F. Wheeler Loomis. Can physics serve two masters? Bulletin of the Atomic Scientists, 6(4):115–120, 127, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopes:1966:SDV

[Lop66] J. Leite Lopes. Science for development — a view from Latin America. Bulletin of the Atomic Scientists, 22(7):7–11, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopes:1968:SMD

[Lop68] J. Leite Lopes. Science for man: The development gap. Bulletin of the Atomic Scientists, 24(10):14–17, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopes:1976:SPH

[Lop76] J. Leite Lopes. A scientist’s plea for human rights in South Africa. Bulletin of the Atomic Scientists, 32(10):3, December 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lev77a].

Lopes:1978:ADN

[Lop78] J. Leite Lopes. Atoms in the developing nations: Brazilian scientists enter the debate. Bulletin of the Atomic Scientists, 34(4): 31–34, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopez:1990:WGW

[Lop90] George A. Lopez. Why the generals wage war on the people. Bulletin of the Atomic Scientists, 46(4):30–33, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopez:1991:GWC

[Lop91] George A. Lopez. The Gulf War: Not so clean. Bulletin of the Atomic Scientists, 47(7):30–35, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HPR91].

Lorenz:1957:LES

[Lor57] Roger S. Lorenz. Letter to the Editor: Sublimate the war impulse? Bulletin of the Atomic Scientists, 13(8):313, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lorenz:1970:KMO

[Lor70] Konrad Lorenz. On killing members of one’s own species. Bulletin of the Atomic Scientists, 26(8):2–5, 51–56, October 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ryb71].

Loran:1974:LWN

[Lor74] William Loran. Letter: We need help. Bulletin of the Atomic Scientists, 30(8):2–3, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Loraine:1977:TDT

[Lor77] John A. Loraine. Time for doctors to take a stand on nuclear proliferation. Bulletin of the Atomic Scientists, 33(8):6–7, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel76a, Fre78].

Loraine:1981:EHH

[Lor81] John A. Loraine. Energy, the hinge of history. Bulletin of the Atomic Scientists, 37(3):19–23, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lorber:1992:TMA

[Lor92] Azriel Lorber. Tactical missiles: Anyone can play. Bulletin of the Atomic Scientists, 48(2):38–40, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lortie:1999:WW

[Lor99] Bret Lortie. Web watch. Bulletin of the Atomic Scientists, 55 (4):13, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lortie:2000:TTW

[Lor00a] Bret Lortie. And then they went home. Bulletin of the Atomic Scientists, 56(3):7, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lortie:2000:WIG

[Lor00b] Bret Lortie. Where’s it gone? The peace movement at the turn of the century. Bulletin of the Atomic Scientists, 56(2):52–56, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
2/52.

Lortie:2001:DIY

[Lor01a] Bret Lortie. A do-it-yourself SIOP. Bulletin of the Atomic Scientists, 57(4):22–29, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/4/22.

Lortie:2001:ICO

[Lor01b] Bret Lortie. It came from Ohio. Bulletin of the Atomic Scientists, 57(2):8, March/April 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lortie:2001:NTS

[Lor01c] Bret Lortie. Nuclear testing: Setting the scene. Bulletin of the Atomic Scientists, 57(2):50–52, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/50.

Lortie:2001:RRT

[Lor01d] Bret Lortie. Russia: Radiation on the Tom. Bulletin of the Atomic Scientists, 57(3):19–21, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/19.

Lortie:2002:NTB

[Lor02] Bret Lortie. Nuclear transport: The big “What–If”. Bulletin of the Atomic Scientists, 58(1):16–18, January/February 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/1/16.

Scientists:1954:SPP

[Los54] Los Alamos Branch of the Federation of Atomic Scientists. A six-point program for balanced security. Bulletin of the Atomic Scientists, 10(7):299, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LACFAS:1957:PNW

[Los57] Los Alamos Chapter of the Federation of American Scientists. Putting nuclear weapons testing in perspective. Bulletin of the Atomic Scientists, 13(7):266, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Loutit:1951:B

[Lou51] J. F. Loutit. Biology. Bulletin of the Atomic Scientists, 7(7–8): 233, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Louria:1990:RHZ

[Lou90] Donald B. Louria. Radiation and health: Zapping the food supply. Bulletin of the Atomic Scientists, 46(7):34–36, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [PGL91].

Lovell:1963:ICS

[Lov63] Bernard Lovell. International cooperation: Soviet aims in astronomy and space research. Bulletin of the Atomic Scientists, 19(8): 36–39, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovell:1966:SMM

[Lov66] Sir Bernard Lovell. [Science and a manned Moon landing]. Bulletin of the Atomic Scientists, 22(9):35, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovell:1968:PS

[Lov68] Bernard Lovell. The pollution of space. Bulletin of the Atomic Scientists, 24(10):42–45, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovell:1969:MMU

[Lov69] Sir Bernard Lovell. Man moves into the universe. Bulletin of the Atomic Scientists, 25(7):4–7, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovins:1973:CAF

[Lov73] Amory B. Lovins. The case against the fast breeder reactor: An anti-nuclear establishment view. Bulletin of the Atomic Scientists, 29(3):29–35, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovins:1974:LR

[Lov74a] Amory B. Lovins. Letter: Response. Bulletin of the Atomic Scientists, 30(9):2–3, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lov74b, Bar74].

Lovins:1974:WESa

[Lov74b] Amory B. Lovins. World energy strategies: Facts, issues and opinions for policies that transcend today’s ‘ad-hocracy’. Bulletin of the Atomic Scientists, 30(5):14–32, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DPL75].

Lovins:1974:WESb

[Lov74c] Amory B. Lovins. World energy strategies/part 2: The case for long-term planning. Bulletin of the Atomic Scientists, 30(6):38– 50, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bar74, DPL75] and response [Lov74a].

Lovins:1978:LC

[Lov78a] Amory B. Lovins. Letter: A convergence? Bulletin of the Atomic Scientists, 34(5):4, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei78f].

Lovins:1978:LFM

[Lov78b] Amory B. Lovins. Letter: Fission not the method. Bulletin of the Atomic Scientists, 34(9):62–63, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovins:1979:TCP

[Lov79] Amory B. Lovins. Thorium cycles and proliferation. Bulletin of the Atomic Scientists, 35(2):16–22, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [vHWL79, DLF79].

Lovins:1986:LIN

[Lov86] Amory B. Lovins. Letter: Iraq’s nuclear intentions. Bulletin of the Atomic Scientists, 42(8):55–56, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil86b].

Lovins:1989:LCF

[Lov89] Amory B. Lovins. Letter: Cold fusion confusion. Bulletin of the Atomic Scientists, 45(5):46–47, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lovering:2000:LEP

[Lov00] Daniel Lovering. Laos: Exploding the past. Bulletin of the Atomic Scientists, 56(5):28–34, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/28.

Lovins:2013:EUC

[Lov13] Amory B. Lovins. The economics of a US civilian nuclear phaseout. Bulletin of the Atomic Scientists, 69(2):44–65, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/2/44.

Lowe:1965:CSA

[Low65] George E. Lowe. Congress and scientific advice. Bulletin of the Atomic Scientists, 21(10):39–42, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lowe:1966:CA

[Low66] George E. Lowe. The Camelot affair. Bulletin of the Atomic Scientists, 22(5):44–48, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lowry:1976:LTS

[Low76] Ritchie P. Lowry. Letter: Too soft on secrecy. Bulletin of the Atomic Scientists, 32(4):2, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Chu76].

Lowry:1985:LUK

[Low85] David Lowry. Letter: United Kingdom and nonproliferation. Bulletin of the Atomic Scientists, 41(8):62–64, September 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HA85].

Lowen:1993:BRT

[Low93] Rebecca Lowen. Book review: Thinking in the bathtub: Genius in the Shadows, by William Lanouette with Bela Silard. Bulletin of the Atomic Scientists, 49(5):51–52, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lowery:1998:BCI

[Low98] Todd Lowery. Bulletin: Call it a MIRACL. Bulletin of the Atomic Scientists, 54(1):5–6, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Longstreth:1985:USP

[LP85a] Thomas K. Longstreth and John E. Pike. U.S., Soviet programs threaten ABM Treaty. Bulletin of the Atomic Scientists, 41(4): 11–15, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LopesdaSilva:1985:LSN

[LP85b] Carlota Lopes da Silva and William Proxmire. Letter: Superpowers’ nonproliferation responsibilities. Bulletin of the Atomic Scientists, 41(6):53–54, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pro85].

Lewis:1991:SSI

[LP91] George Lewis and Theodore Postel. START: SLCMs — ignored, then stored. Bulletin of the Atomic Scientists, 47(9):26– 28, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindsay-Poland:1992:BRH

[LP92] John Lindsay-Poland. Book review: Hooked on drug wars: Cocaine Politics: Drugs, Armies and the CIA in Central America, by Peter Dale Scott and Jonathan Marshall. Bulletin of the Atomic Scientists, 48(3):45–46, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1997:PBI

[LP97] George Lewis and Theodore Postol. Portrait of a bad idea. Bulletin of the Atomic Scientists, 53(4):18–25, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WLP97].

Lewis:2008:EMD

[LP08] George N. Lewis and Theodore A. Postol. The European missile defense folly. Bulletin of the Atomic Scientists, 64(2):32–39, May/ June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/64/2/32.

Lewis:2010:HUS

[LP10] George N. Lewis and Theodore A. Postol. How US strategic antimissile defense could be made to work. Bulletin of the Atomic Scientists, 66(6):8–24, November/December 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/6/8.

Livingston:1954:OLP

[LPT^{+}54] M. Stanley Livingston, Ernest C. Pollard, Lewis Tonks, Watson Davis, James Franck, Oscar K. Rice, W. A. Higinbotham, David L. Hill, and John S. Toll. Open letter to President Eisenhower. Bulletin of the Atomic Scientists, 10(6):255, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lindop:1971:RPE

[LR71] Patricia J. Lindop and J. Rotblat. Radiation pollution of the environment. Bulletin of the Atomic Scientists, 27(7):17–24, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

LaRocque:1979:SRP

[LR79] Gene R. La Rocque and George W. Rathjens. Special report: Panel II: The first nuclear war conference. Bulletin of the Atomic Scientists, 35(5):24, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Logan:1991:LEH

[LRD^{+}91] Jonothan L. Logan, Helmut Rechenberg, Max Dresden, A. Van Der Ziel, and Mark Walker. Letters to the Editor: Heisenberg, Goudsmit and the German “A-bomb”. Physics Today, 44(5): 13, 15, 90–92, 94–96, May 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v44/i5/p13_s1. Let-

ters in response to [Wal90b], including Walker’s rebuttal.

Laughlin:1958:RMP

[LS58] John S. Laughlin and Robert S. Sherman. Radiation and man: Peacetime radiation: Radiation exposure incidental to medical practice. Bulletin of the Atomic Scientists, 14(1):41–43, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martyl:1959:SAI

[LS59] Martyl Langsdorf and Cyril Stanley Smith. Science and art: Introduction. Bulletin of the Atomic Scientists, 15(2):50–51, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1965:LED

[LS65] David K. Lewis and Jeremy J. Stone. Letters to the Editor: Defense and disarmament. Bulletin of the Atomic Scientists, 21 (1):29–31, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sto64].

Landy:1992:DMC

[LS92] Joanne Landy and Jennifer Scarlot. Democratic movements can force disarmament. Bulletin of the Atomic Scientists, 48(4):29– 31, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lanouette:1994:GSB

[LS94] William Lanouette and Bela A. Silard. Genius in the shadows: a biography of Leo Szilard: the man behind the bomb. University of Chicago Press, Chicago, IL, USA and London, UK, 1994. ISBN 0-226-46888-7 (paperback). xix + 587 pp. LCCN QC16.S95

L36 1994. URL http://www.loc.gov/catdir/description/
uchi053/94012738.html. Foreword by Jonas Salk.

Langsdorf:2020:SA

[LS20] Martyl Langsdorf and Cyril Stanley Smith. 1959: Science and art. Bulletin of the Atomic Scientists, 76(6):363–365, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847483.

Lecht:1974:LRP

[LSB74] Jane G. Lecht, Pushpa N. Schwartz, and Norman Brown. Letter: Renaissance playgirl. Bulletin of the Atomic Scientists, 30(7):2– 3, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day74c].

Lopez:1995:GT

[LSP95a] George A. Lopez, Jackie G. Smith, and Ron Pagnucco. The global tide. Bulletin of the Atomic Scientists, 51(4):33–39, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lopez:1995:TMD

[LSP95b] George A. Lopez, Jackie G. Smith, and Ron Pagnucco. “Tizmoes” do it best. Bulletin of the Atomic Scientists, 51(4):37, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lieggi:2010:TCS

[LST10] Stephanie Lieggi, Robert Shaw, and Masako Toki. Taking control: Stopping North Korean WMD-related procurement. Bulletin of the Atomic Scientists, 66(5):21–34, September/October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/5/21.

Lerner:1970:IWP

[LT70] Daniel Lerner and Albert H. Teich. Internationalism and world politics among CERN scientists. Bulletin of the Atomic Scientists, 26(2):4–10, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Larin:1999:SPL

[LT99a] Vladislav Larin and Eugeny Tar. Soviet PNEs: a legacy of contamination. Bulletin of the Atomic Scientists, 55(3):18–20, May/ June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/55/3/18.

Latin:1999:LC

[LT99b] Vladislav Latin and Eugeny Tar. A legacy of contamination. Bulletin of the Atomic Scientists, 55(3):18, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lubkemeier:1990:NIC

[Lüb90] Eckhard Lübkemeier. NATO’s identity crisis. Bulletin of the Atomic Scientists, 46(8):30–33, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lucey:1982:SUR

[Luc82] Paul A. Lucey. The Soviet Union: Religious institutions and practices. Bulletin of the Atomic Scientists, 38(5):31–36, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lucas:1990:GUS

[Luc90a] Michael R. Lucas. Get used to saying ‘CSCE’. Bulletin of the Atomic Scientists, 46(9):32–34, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Luckham:1990:PUP

[Luc90b] Robin Luckham. Perspective: U.S. projects power, paradox in Gulf. Bulletin of the Atomic Scientists, 46(10):10–11, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lueders:1996:ESL

[Lue96] Bill Lueders. An Enemy of the State: the Life of Erwin Knoll. Common Courage Press, Monroe, ME, USA, 1996. ISBN 1-56751-098-1, 1-56751-099-X. v + 297 pp. LCCN PN4874.K62 L84 1996.

Lugar:2000:RLB

[Lug00] Richard Lugar. Richard Lugar: Better safe. … Bulletin of the Atomic Scientists, 56(1):44–46, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/1/44.

Lugar:2005:ACR

[Lug05] Richard G. Lugar. Article commentary: Redefining the threat. Bulletin of the Atomic Scientists, 61(5):20–21, September/ October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
5/20.

Lumpe:1994:LSJ

[Lum94a] Lora Lumpe. A license to steal jobs. Bulletin of the Atomic Scientists, 50(5):32, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lumpe:1994:SDS

[Lum94b] Lora Lumpe. Sweet deals, stolen jobs. Bulletin of the Atomic Scientists, 50(5):30–35, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lumpe:1996:CG

[Lum96] Lora Lumpe. Costly giveaways. Bulletin of the Atomic Scientists, 52(5):30–38, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lumpe:1999:LP

[Lum99] Lora Lumpe. The leader of the pack. Bulletin of the Atomic Scientists, 55(1):27–33, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/27.

Lundbergh:1959:SAE

[Lun59] Holger Lundbergh. Sweden’s atomic energy program. Bulletin of the Atomic Scientists, 15(5):219–220, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lundberg:1965:SA

[Lun65] B. K. O. Lundberg. The supersonic adventure. Bulletin of the Atomic Scientists, 21(2):29–33, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lundin:1979:SCW

[Lun79] S. J. Lundin. From Sweden: Chemical weapons: too late for disarmament? Bulletin of the Atomic Scientists, 35(10):33–37, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lunn:1982:INM

[Lun82] Simon Lunn. At issue: nuclear modernization in Europe. Bulletin of the Atomic Scientists, 38(7):17–23, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Luria:1977:GS

[Lur77] S. E. Luria. The goals of science. Bulletin of the Atomic Scientists, 33(5):28–33, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Luten:1973:CCN

[Lut73] Daniel B. Luten. Communications: A chap named Huxley. Bulletin of the Atomic Scientists, 29(7):55, September 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fre72b].

Luttwak:1989:ACC

[Lut89] Edward N. Luttwak. Arms control criteria: Close, but no cigar. Bulletin of the Atomic Scientists, 45(4):35, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:2018:LBM

[LvH18] George Lewis and Frank von Hippel. Limitations on ballistic missile defense — past and possibly future. Bulletin of the Atomic Scientists, 74(4):199–209, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2018.
1486575.

Livingston:1961:FPA

[LW61] M. Stanley Livingston and Hugh C. Wolfe. FAS position on arms control. Bulletin of the Atomic Scientists, 17(4):137–139, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lewis:1969:NSV

[LW69] John W. Lewis and Jayne S. Werner. The “New Stage” in Vietnam. Bulletin of the Atomic Scientists, 25(1):21–26, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lwoff:1960:PSI

[Lwo60] Andre Lwoff. The possibility of selecting infectious disease agents. Bulletin of the Atomic Scientists, 16(6):243–244, June 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyford:1961:LED

[Lyf61] Cabot Lyford. Letter to the Editor: The danger of waiting. Bulletin of the Atomic Scientists, 17(1):0, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyman:1996:WPJ

[Lym96] Edwin S. Lyman. Weapons plutonium: Just can it. Bulletin of the Atomic Scientists, 52(6):48–52, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gar97b].

Lyman:1997:LCP

[Lym97] Edwin Lyman. Letter: Clarification: Plutonium: Bury the stuff. Bulletin of the Atomic Scientists, 53(3):60–61, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LL97].

Lyman:2008:CNP

[Lym08] Edwin S. Lyman. Can nuclear plants be safer? Bulletin of the Atomic Scientists, 64(4):34–37, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/4/34.

Lyman:2011:SOT

[Lym11] Edwin S. Lyman. Surviving the one-two nuclear punch: Assessing risk and policy in a post-Fukushima world. Bulletin of the Atomic Scientists, 67(5):47–54, September/October 2011. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/67/5/47.

Lynch:1969:BBA

[Lyn69] William S. Lynch. Books: The American University, by Jacques Barzun; Crisis at Columbia; Up Against the Ivy Wall, by Jerry L. Avorn and others; Revolution at Berkeley, edited by Michael V. Miller and Susan Gilmore; Democracy and the Student Left, by George F. Kennan. Bulletin of the Atomic Scientists, 25(8):

38–41, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lynch:1978:BRB

[Lyn78] Virginia Lynch. Book reviews: The Poison that Fell from the Sky, by John G. Fuller; Heart Beat, by Eugene Dong and Spyros Andreopoulos. Bulletin of the Atomic Scientists, 34(10):44–45, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lynn:1984:CFB

[Lyn84] Richard Lynn. Commentary: Flaw in build-down. Bulletin of the Atomic Scientists, 40(10):34, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BW84].

Lynch:1986:PCA

[Lyn86] John Lynch. The Pentagon’s community aid program. Bulletin of the Atomic Scientists, 42(8):34–37, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lynch:1988:RSF

[Lyn88] Allen Lynch. The restructuring of Soviet foreign policy. Bulletin of the Atomic Scientists, 44(2):40–43, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyons:1963:MM

[Lyo63] Gene M. Lyons. The military mind. Bulletin of the Atomic Scientists, 19(9):19–22, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyttle:1979:BRB

[Lyt79] Bradford Lyttle. Book reviews: 1982, by Norman Alcock, William Eckhardt, Anita Kemp, Gernot Köhler, Hanna Newcombe, Valerie Restivo, and Arnold Simoni. With a foreword by Jan Tinbergen. Bulletin of the Atomic Scientists, 35(3):67, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Lyttle:1985:LDM

[Lyt85] Bradford Lyttle. Letter: Don’t move the clock back. Bulletin of the Atomic Scientists, 41(3):61–62, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Phi84b].

M:1984:BRB

[M.84] S. M. Book review: Science and Technology in India, by A. Rahman. Bulletin of the Atomic Scientists, 40(9):53, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

M:1985:BNB

[M.85] S. M. Book note: The Nuclear Weapons Industry, by Kenneth S. Shaw. Bulletin of the Atomic Scientists, 41(1):53, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maruf:1952:LEE

[Ma’52] H. Farid Ma’ruf. Letter to the Editor: [The ethical problems of science]. Bulletin of the Atomic Scientists, 8(1):33, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lon51, Wyl51, Tel51a].

Mills:1983:CEI

[MA83] Mark P. Mills and John Abbotts. Commentary: Electricity and industry. Bulletin of the Atomic Scientists, 39(2):50–51, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Abb82] and erratum [Ano83l].

Marchant:2017:SLN

[MA17] Gary E. Marchant and Brad Allenby. Soft law: New tools for governing emerging technologies. Bulletin of the Atomic Scientists, 73(2):108–114, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1288447.

Mabogunje:1969:ADA

[Mab69] Akin L. Mabogunje. Agricultural development in Africa. Bulletin of the Atomic Scientists, 25(4):21–23, 48, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mabe:1993:RPR

[Mab93] William Mabe. Report: Proliferation: Retrofit Russian research reactors. Bulletin of the Atomic Scientists, 49(3):9–11, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacArthur:1951:GMW

[Mac51] General Douglas MacArthur. General MacArthur: War is obsolete. Bulletin of the Atomic Scientists, 7(6):167–168, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacDougal:1954:LEH

[Mac54a] P. S. MacDougal. Letter to the Editor: H-bomb tests in the Pacific and the firing of Oppenheimer. Bulletin of the Atomic Scientists, 10(8):335, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacLeish:1954:OCL

[Mac54b] Archibald MacLeish. Our capacity to live. Bulletin of the Atomic Scientists, 10(2):36, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Macmillan:1957:HA

[Mac57] Harold Macmillan. How’s that again? Bulletin of the Atomic Scientists, 13(6):194, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maccoby:1961:SPD

[Mac61] Michael Maccoby. Social psychology of deterrence. Bulletin of the Atomic Scientists, 17(7):278–281, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mackintosh:1962:BBD

[Mac62] Malcolm Mackintosh. Books: Civil Defense in the Soviet Union, by Leon Gouré. Bulletin of the Atomic Scientists, 18(6):38–39, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mackintosh:1965:MAS

[Mac65a] Malcolm Mackintosh. The military aspects of the Sino–Soviet dispute. Bulletin of the Atomic Scientists, 21(8):14–17, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacLane:1965:CWB

[Mac65b] Saunders MacLane. Can we buy quality in science? Bulletin of the Atomic Scientists, 21(9):6–11, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacRae:1966:CSP

[Mac66] Duncan MacRae, Jr. Careers, science, and politics. Bulletin of the Atomic Scientists, 22(2):26–28, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacDonald:1967:SSP

[Mac67] Gordon J. F. MacDonald. Science and space policy: How does it get planned? Bulletin of the Atomic Scientists, 23(5):2–9, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacDonald:1968:SPR

[Mac68a] Gordon J. F. MacDonald. Science and politics of rainmaking. Bulletin of the Atomic Scientists, 24(8):8–14, October 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ney69].

Mackenzie:1968:PT

[Mac68b] William George Mackenzie. Panjandrums of technology. Bulletin of the Atomic Scientists, 24(9):40, November 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab68e].

MacBride:1977:NMN

[Mac77a] Sean MacBride. A new morality for a new world. Bulletin of the Atomic Scientists, 33(7):22–23, September 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Cultures 3(4) 1976. See comment [Hue77].

Macioti:1977:SPV

[Mac77b] Manfredo Macioti. A science policy visit to Israel. Bulletin of the Atomic Scientists, 33(3):10–21, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mackintosh:1977:LAR

[Mac77c] A. R. Mackintosh. Letter: Acceptable risks. Bulletin of the Atomic Scientists, 33(10):5, 8–9, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol77].

MacBride:1981:NNN

[Mac81a] Sean MacBride. Namibia, Namibia, Namibia. Bulletin of the Atomic Scientists, 37(6):22–24, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mack:1981:CSF

[Mac81b] John E. Mack. Commentary from San Francisco: Psychosocial effects of the nuclear arms race. Bulletin of the Atomic Scientists, 37(4):18–23, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maclaren:1981:CB

[Mac81c] Peirs Maclaren. Commentary: On the beach. Bulletin of the Atomic Scientists, 37(9):63, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Nor82].

MacDonald:1983:CHR

[Mac83] Wally MacDonald. Commentary: On ‘Human Rights Imperialism’. Bulletin of the Atomic Scientists, 39(9):63, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Von83].

MacDonald:1984:CSA

[Mac84a] Wally MacDonald. Commentary: Slaves to animosity. Bulletin of the Atomic Scientists, 40(9):57, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bal84b].

Mack:1984:BRB

[Mac84b] John E. Mack. Book review: Against the State of Nuclear Terror, by Joel Kovel. Bulletin of the Atomic Scientists, 40(5):44, 46–47, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacDougall:1986:LPC

[Mac86a] John MacDougall. Letter: Pressuring Congress. Bulletin of the Atomic Scientists, 42(4):54–55, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacDougall:1986:LSE

[Mac86b] John MacDougall. Letter: Soft energy path. Bulletin of the Atomic Scientists, 42(10):52, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro86d].

Mackenzie:1986:MAA

[Mac86c] Donald Mackenzie. Missile accuracy — an arms-control opportunity. Bulletin of the Atomic Scientists, 42(6):11–16, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mack:1988:BCC

[Mac88] Andrew Mack. Benefits of cruise control. Bulletin of the Atomic Scientists, 44(5):7–9, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fel88].

MacDougall:1989:LFG

[Mac89] John MacDougall. Letter: Flourishing grassroots. Bulletin of the Atomic Scientists, 45(9):50, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Boy89a, SM89a].

MacDougall:1991:LEG

[Mac91] John MacDougall. Letter: End global arms sales. Bulletin of the Atomic Scientists, 47(7):47, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sol91].

MacDougall:1992:LGC

[Mac92] John MacDougall. Letter: Grassroots can say no to Pentagon. Bulletin of the Atomic Scientists, 48(9):46, November 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MacKenzie:1996:GOM

[Mac96] Donale MacKenzie. Guest opinion: Moving toward disinvention. Bulletin of the Atomic Scientists, 52(5):4, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mack:1997:CK

[Mac97a] Andrew Mack. “Common knowledge”. Bulletin of the Atomic Scientists, 53(4):51, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mack:1997:PP

[Mac97b] Andrew Mack. Potential, not proliferation. Bulletin of the Atomic Scientists, 53(4):48–53, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Macfarlane:1998:LJD

[Mac98] Allison Macfarlane. Letter: Jackass decision? Bulletin of the Atomic Scientists, 54(3):4–5, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98].

MacFarlane:2006:SS

[Mac06] Allison MacFarlane. Stuck on a solution. Bulletin of the Atomic Scientists, 62(3):46–52, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/3/46.

Macfarlane:2011:DYK

[Mac11] Allison Macfarlane. It’s 2050: Do you know where your nuclear waste is? Bulletin of the Atomic Scientists, 67(4):30–36, July/ August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
4/30.

Macfarlane:2022:RAF

[Mac22] Douglas R. Macfarlane. Renewable ammonia: The future of fuels? Bulletin of the Atomic Scientists, 79(1):14–16, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2155005.

Maddox:1963:LWK

[Mad63] John Maddox. Is the literature worth keeping? Bulletin of the Atomic Scientists, 19(9):14–16, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maddox:1979:BRB

[Mad79] John Maddox. Book reviews: The Human Future Revisited, by Harrison Brown. Bulletin of the Atomic Scientists, 35(3):67, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maerli:2001:NSH

[Mae01] Morten Bremer Maerli. The Navy says: Hands off my stash. Bulletin of the Atomic Scientists, 57(5):24–25, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/5/24.

Maguire:1973:CTF

[Mag73] Catherine Maguire. Communications: Tomorrow’s fossils. Bulletin of the Atomic Scientists, 29(1):3, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fra72].

Magat:1981:BRF

[Mag81] Richard Magat. Book review: A forgotten novel remembered: The Accident, by Dexter Masters. Bulletin of the Atomic Scientists, 37(1):49–50, 52–53, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WM82].

MagoAmilcar:1985:LSA

[Mag85] Hanno MagoAmilcar. Letter: South African bomb. Bulletin of the Atomic Scientists, 41(10):54–55, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Magraw:1988:TCB

[Mag88a] Katherine Magraw. Teller and the “clean bomb” episode. Bulletin of the Atomic Scientists, 44(4):32, 34–37, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Magraw:1988:WRS

[Mag88b] Katherine Magraw. Washington report: Sniffing out SLCMs; showdown on the slippery slope; hairy-chested Democrats; Midgetman plan falls short. Bulletin of the Atomic Scientists, 44(7):3–5, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Magno:2016:PAN

[Mag16] Paul Magno. The Plowshares anti-nuclear movement at 35: a next generation? Bulletin of the Atomic Scientists, 72(2):85– 88, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mahalanobis:1959:TPCb

[Mah59a] P. C. Mahalanobis. Third Pugwash Conference: Appendix: Economic development of Afro–Asian countries. Bulletin of the Atomic Scientists, 15(1):46, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mahalanobis:1959:TPCa

[Mah59b] P. C. Mahalanobis. Third Pugwash Conference: Industrialization of underdeveloped countries — a means to peace. Bulletin of the Atomic Scientists, 15(1):12–17, 46, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also appendix [Mah59a].

Makhijani:1976:SER

[Mak76] Arjun Makhijani. Solar energy and rural development for Third World. Bulletin of the Atomic Scientists, 32(6):14–24, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1990:CSF

[Mak90] Arjun Makhijani. Common security is far off. Bulletin of the Atomic Scientists, 46(4):41–42, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1994:EEG

[Mak94a] Arjun Makhijani. Energy enters guilty plea. Bulletin of the Atomic Scientists, 50(2):18–20, 25–28, March/April 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano94p].

Makhijani:1994:L

[Mak94b] Arjun Makhijani. Let’s not. Bulletin of the Atomic Scientists, 50 (3):44–45, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1995:AT

[Mak95a] Arjun Makhijani. “Always” the target? Bulletin of the Atomic Scientists, 51(3):23–27, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1995:GOO

[Mak95b] Arjun Makhijani. Guest opinion: Open the files, please. Bulletin of the Atomic Scientists, 51(1):4, January/February 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1997:IO

[Mak97] Arjun Makhijani. India’s options. Bulletin of the Atomic Scientists, 53(2):51–53, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1998:LL

[Mak98] Arjun Makhijani. A legacy lost. Bulletin of the Atomic Scientists, 54(4):53–56, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:2001:BP

[Mak01] Arjun Makhijani. The burden of proof. Bulletin of the Atomic Scientists, 57(4):49–54, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/4/49.

Makhijani:2003:NTF

[Mak03] Arjun Makhijani. Nuclear targeting: The first 60 years. Bulletin of the Atomic Scientists, 59(3):60–65, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/3/60.

Maksymiuk:2006:EED

[Mak06] Jan Maksymiuk. Eastern Europe: Diplomatic dustup. Bulletin of the Atomic Scientists, 62(1):16–17, January/February 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/1/16.

Makhijani:2024:EFE

[Mak24] Arjun Makhijani. The entanglement of fusion energy research and bombs. Bulletin of the Atomic Scientists, 80(6):352–357, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2416833.

Malina:1950:ICS

[Mal50] Frank J. Malina. International cooperation in science: The work of UNESCO. Bulletin of the Atomic Scientists, 6(4):121–125, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Malik:1954:SRT

[Mal54] Charles Malik. Some reflections on technical and economic assistance. Bulletin of the Atomic Scientists, 10(3):93–96, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Malmlow:1963:APS

[Mal63] E. G. Malmlöw. Atomic power in Sweden. Bulletin of the Atomic Scientists, 19(2):44–46, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Malamud:1965:LEK

[Mal65] Herbert Malamud. Letter to the Editor: “Kick the scientist”. Bulletin of the Atomic Scientists, 21(5):34–35, May 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Malone:1979:PUF

[Mal79] Thomas F. Malone. Perspective: UNCSTD: a first step in a long journey. Bulletin of the Atomic Scientists, 35(9):1, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mallow:1981:CC

[Mal81] Jeffry V. Mallow. Commentary: ‘Of course’. Bulletin of the Atomic Scientists, 37(6):64, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alf81].

Malone:1985:ISN

[Mal85] Thomas F. Malone. International scientists on nuclear winter. Bulletin of the Atomic Scientists, 41(11):52–54, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Malik:2014:WWR

[Mal14] Salma Malik. Women and weapons: Redressing the gender gap: a Pakistani response. Bulletin of the Atomic Scientists, 70(5):12– 16, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manely:1949:AL

[Man49a] J. H. Manely. The Los Alamos Laboratory. Bulletin of the Atomic Scientists, 5(4):101–105, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manley:1949:ASL

[Man49b] J. H. Manley. The Los Alamos Scientific Laboratory. Bulletin of the Atomic Scientists, 5(4):101–103, 105, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mann:1957:LEV

[Man57] Lawrence Mann. Letter to the Editor: The value presuppositions of science. Bulletin of the Atomic Scientists, 13(6):231–232, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bur57].

Manley:1959:SC

[Man59] John H. Manley. Science and crisis. Bulletin of the Atomic Scientists, 15(3):114–117, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mann:1964:WSO

[Man64] Georg Mann. Why not settle for one good culture? Bulletin of the Atomic Scientists, 20(8):30–31, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar63, Bar64a].

Mandel:1966:LEI

[Man66a] Loring Mandel. Letter to the Editor: On involvement. Bulletin of the Atomic Scientists, 22(4):31, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mangone:1966:WDU

[Man66b] Gerard J. Mangone. What does the U.S. want from the U.N.? Bulletin of the Atomic Scientists, 22(1):8–12, January 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manheim:1972:CCE

[Man72] Frank T. Manheim. Communications: 18th Century Enlightenment. Bulletin of the Atomic Scientists, 28(10):2, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi72b].

Manley:1974:AOT

[Man74] J. H. Manley. “All in our time”: Assembling the wartime labs. Bulletin of the Atomic Scientists, 30(5):42–48, May 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mandelbaum:1977:NEC

[Man77] Michael Mandelbaum. A nuclear exporters cartel: It could not halt proliferation permanently, but it might give the world a few more years to find a formula. Bulletin of the Atomic Scientists, 33(1):42–50, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mandelbaum:1979:DS

[Man79] Michael Mandelbaum. In defense of SALT. Bulletin of the Atomic Scientists, 35(1):15–21, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manganyi:1980:NSD

[Man80] N. C. Manganyi. The North–South dialogue: who writes the script? Bulletin of the Atomic Scientists, 36(10):12–14, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mandelbaum:1984:ANA

[Man84a] David C. Mandelbaum. Anthropology for the nuclear age. Bulletin of the Atomic Scientists, 40(6):11–15, June/July 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manley:1984:CMH

[Man84b] John H. Manley. Commentary: In the matter of the H-bomb. Bulletin of the Atomic Scientists, 40(1):52–53, January 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manoff:1984:AMN

[Man84c] Robert Karl Manoff. Agenda 1984: The media: nuclear secrecy vs. democracy. Bulletin of the Atomic Scientists, 40(1):26–29, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mandelbaum:1985:FUS

[Man85] Michael Mandelbaum. Futures: Uncertainty of the status quo. Bulletin of the Atomic Scientists, 41(7):131–135, August 1985.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mann:1986:NSW

[Man86] Paul Mann. The NASA story we missed. Bulletin of the Atomic Scientists, 42(9):18, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manley:1987:RTH

[Man87] John Manley. Return technology to human hands. Bulletin of the Atomic Scientists, 43(10):7–8, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mandelbaum:1989:ACC

[Man89] Michael Mandelbaum. Arms control criteria: Facts, not pacts. Bulletin of the Atomic Scientists, 45(4):44–45, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Manning:1995:AVB

[Man95] Mary Manning. Atomic vets battle time. Bulletin of the Atomic Scientists, 51(1):54–60, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mangano:2004:TMI

[Man04] Joseph Mangano. Three Mile Island: Health study meltdown. Bulletin of the Atomic Scientists, 60(5):30–35, September/ October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/60/
5/30.

Mann:2015:SSH

[Man15] Michael E. Mann. The Serengeti strategy: How special interests try to intimidate scientists, and how best to fight back. Bulletin of the Atomic Scientists, 71(1):33–45, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mann:2021:PNH

[Man21] Michael Mann. The President needs to hit the ground running on climate. Bulletin of the Atomic Scientists, 77(1):21–23, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1860325.

Mann:2023:BEC

[Man23] Michael E. Mann. Book excerpt — catastrophic climate change: Lessons from the dinosaurs. Bulletin of the Atomic Scientists, 79 (6):353–359, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2266936.

Marschak:1946:EAA

[Mar46a] J. Marschak. The economic aspects of atomic power. Bulletin of the Atomic Scientists, 2(5–6):8–9, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marschak:1946:EAP

[Mar46b] J. Marschak. The economics of atomic power. Bulletin of the Atomic Scientists of Chicago, 1(5):3, 11, February 15, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maritain:1948:SC

[Mar48a] Jacques Maritain. The scientist and the community. Bulletin of the Atomic Scientists, 4(12):367, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1948:AEA

[Mar48b] Herbert S. Marks. The Atomic Energy Act: Public administration without public debate. Bulletin of the Atomic Scientists, 4 (9):271–276, 288, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshak:1948:PSN

[Mar48c] Robert E. Marshak. Present state of the UN negotiations on atomic energy. Bulletin of the Atomic Scientists, 4(1):2, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshall:1948:PC

[Mar48d] George C. Marshall. Policy at the crossroads. Bulletin of the Atomic Scientists, 4(9):261–262, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1948:LEH

[Mar48e] Abram V. Martin. Letter to the Editors: How to implement Dr. Szilard’s proposal to Mr. Stalin. Bulletin of the Atomic Scientists, 4(3):80, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marinin:1949:ESE

[Mar49a] M. Marinin. An English scientist exposes atomic diplomacy. Bulletin of the Atomic Scientists, 5(2):43, 50, February 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maritain:1949:SMH

[Mar49b] Jacques Maritain. Science, materialism, and the human spirit: Science and ontology. Bulletin of the Atomic Scientists, 5(6–7): 199–200, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1949:WNe

[Mar49c] Anne Wilson Marks. Washington notes. Bulletin of the Atomic Scientists, 5(5):158–160, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1949:WNh

[Mar49d] Anne Wilson Marks. Washington notes. Bulletin of the Atomic Scientists, 5(8–9):218–220, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1949:WNl

[Mar49e] Anne Wilson Marks. Washington notes. Bulletin of the Atomic Scientists, 5(12):327–328, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1949:WNA

[Mar49f] Anne Wilson Marks. Washington notes: AEC under fire. Bulletin of the Atomic Scientists, 5(6–7):164–165, 181, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1949:CA

[Mar49g] Herbert S. Marks. Congress and the atom. Bulletin of the Atomic Scientists, 5(2):44–47, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1951:WN

[Mar51] Anne Wilson Marks. Washington notes. Bulletin of the Atomic Scientists, 7(1):17–18, 22, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1953:WNa

[Mar53a] Anne W. Marks. Washington notes. Bulletin of the Atomic Scientists, 9(3):83–84, 94, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marks:1953:WNb

[Mar53b] Anne W. Marks. Washington notes. Bulletin of the Atomic Scientists, 9(4):100–102, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1953:LEJ

[Mar53c] A. B. Martin. Letter to the Editor: Jacques Monod and the French views on the Rosenberg case. Bulletin of the Atomic Scientists, 9(10):393, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mon53a].

Marder:1954:FMS

[Mar54a] Murrey Marder. The Fort Monmouth story. Bulletin of the Atomic Scientists, 10(1):21–25, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1954:LES

[Mar54b] Walter R. Marseille. Letter to the Editor: Science and military power. Bulletin of the Atomic Scientists, 10(3):113, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Ber53a].

Marseille:1954:LEP

[Mar54c] Walter S. Marseille. Letter to the Editor: Panic, psychology, and the bomb. Bulletin of the Atomic Scientists, 10(4):145, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wyl54, Mer54b].

Marsh:1954:EDB

[Mar54d] Arthur Marsh. [Economics of developing backward areas]. Bulletin of the Atomic Scientists, 10(3):80, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshall:1954:BBC

[Mar54e] John Marshall. Books: The Cyclotron, by W. B. Mann. Bulletin of the Atomic Scientists, 10(6):232, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshall:1954:BBM

[Mar54f] Leona Marshall. Books: Nuclear Moments, by Norman F. Ramsey. Bulletin of the Atomic Scientists, 10(6):233, June 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marine:1955:AP

[Mar55a] Gene Marine. Atoms in the press. Bulletin of the Atomic Scientists, 11(7):250–252, 264, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1955:NS

[Mar55b] Walter W. Marseille. Negotiation from strength? Bulletin of the Atomic Scientists, 11(1):13–18, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marine:1958:BBV

[Mar58a] Gene Marine. Books: The Voyage of the Lucky Dragon, by Ralph E. Lapp, with a foreword by Pearl S. Buck. Bulletin of the Atomic Scientists, 14(6):236–237, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1958:CAP

[Mar58b] Walter W. Marseille. The case against Pugwash: Not war, not peace. Bulletin of the Atomic Scientists, 14(4):140–143, 159, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1958:LEL

[Mar58c] Walter W. Marseille. Letter to the Editor: Lord Russell and Pugwash. Bulletin of the Atomic Scientists, 14(7):281, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rus58a].

Marshak:1958:NSS

[Mar58d] R. E. Marshak. Nature of the Soviet scientific challenge. Bulletin of the Atomic Scientists, 14(2):83–86, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marine:1959:BBG

[Mar59a] Gene Marine. Books: The Great Decision. The Secret History of the Atomic Bomb, by Michael Amrine. Bulletin of the Atomic Scientists, 15(5):221–222, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1959:RCG

[Mar59b] Walter W. Marseille. Reader’s comment: The great delusion. Bulletin of the Atomic Scientists, 15(8):343–346, October 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

March:1960:NRc

[Mar60a] Robert H. March. News roundup. Bulletin of the Atomic Scientists, 16(3):111–113, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

March:1960:NRd

[Mar60b] Robert H. March. News roundup. Bulletin of the Atomic Scientists, 16(4):142–144, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

March:1960:NRe

[Mar60c] Robert H. March. News roundup. Bulletin of the Atomic Scientists, 16(5):190–192, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marine:1961:PHA

[Mar61a] Gene Marine. Politics haunts astronomers at Berkeley. Bulletin of the Atomic Scientists, 17(8):345–346, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marossi:1961:CUC

[Mar61b] Ruth Marossi. Canada’s uranium crisis. Bulletin of the Atomic Scientists, 17(7):281–286, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1961:RCM

[Mar61c] Walter W. Marseille. Reflections and comments: Marseille replies. Bulletin of the Atomic Scientists, 17(7):294, September 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marseille:1961:RET

[Mar61d] Walter W. Marseille. A review essay: On thermonuclear war. Bulletin of the Atomic Scientists, 17(4):157–159, 166, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Kea61, Sto61, Pet61, Mar61c].

Marseille:1962:AAC

[Mar62] Walter W. Marseille. Attacks ambivalence on CD. Bulletin of the Atomic Scientists, 18(1):38, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshak:1963:RSS

[Mar63] Robert E. Marshak. Reexamining the Soviet scientific challenge. Bulletin of the Atomic Scientists, 19(4):12–17, April 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marcuse:1964:WL

[Mar64a] Herbert Marcuse. World without a logos. Bulletin of the Atomic Scientists, 20(1):25–26, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1964:WBC

[Mar64b] Howard Margolis. From Washington: The bomb in China. Bulletin of the Atomic Scientists, 20(10):36–39, December 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1964:RAC

[Mar64c] Howard Margolis. Reports: From Atlantic City: Notes on the Democratic Convention. Bulletin of the Atomic Scientists, 20(8): 39–42, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1964:RSF

[Mar64d] Howard Margolis. Reports: From San Francisco: Notes on the Republican Convention. Bulletin of the Atomic Scientists, 20 (7):31–34, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1964:RWN

[Mar64e] Howard Margolis. Reports: From Washington: Notes on the MLF. Bulletin of the Atomic Scientists, 20(9):28–30, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1964:RWR

[Mar64f] Howard Margolis. Reports: From Washington: R and D on Capitol Hill. Bulletin of the Atomic Scientists, 20(5):33–37, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1964:RWC

[Mar64g] Howard Margolis. Reports: From Washington: The curious case of Krebiozen. Bulletin of the Atomic Scientists, 20(3):29–31, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ivy64].

Margolis:1964:RWP

[Mar64h] Howard Margolis. Reports: From Washington: The politics of fluoridation. Bulletin of the Atomic Scientists, 20(6):38–41, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kor64, Alb64].

Margolis:1964:RWV

[Mar64i] Howard Margolis. Reports: From Washington: Velikovsky rides again. Bulletin of the Atomic Scientists, 20(4):38–40, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marmor:1964:FWV

[Mar64j] Judd Marmor. Forum: War, violence, and human nature. Bulletin of the Atomic Scientists, 20(3):19–22, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:FFV

[Mar65a] Howard Margolis. Forum: Focus on Vietnam: From Washington: The spring lull. Bulletin of the Atomic Scientists, 21(6):26–27, 48–49, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:WMN

[Mar65b] Howard Margolis. From Washington: McNamara’s new budget. Bulletin of the Atomic Scientists, 21(3):54–56, March 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:WNV

[Mar65c] Howard Margolis. From Washington: Notes on Vietnam. Bulletin of the Atomic Scientists, 21(4):2–3, 47–58, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:WAF

[Mar65d] Howard Margolis. From Washington: The Air Force in space and peacekeeping assessments. Bulletin of the Atomic Scientists, 21(8):34–37, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:RNY

[Mar65e] Howard Margolis. Reports: From New York: Diplomatic game at the U.N. Bulletin of the Atomic Scientists, 21(1):40–44, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:RWNa

[Mar65f] Howard Margolis. Reports: From Washington: Notes on defense matters. Bulletin of the Atomic Scientists, 21(7):36–38, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:RWNb

[Mar65g] Howard Margolis. Reports: From Washington: Notes on gas and disarmament. Bulletin of the Atomic Scientists, 21(9):30– 32, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1965:RWS

[Mar65h] Howard Margolis. Reports: From Washington: Some problems in Vietnam. Bulletin of the Atomic Scientists, 21(5):40–43, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marx:1965:REY

[Mar65i] Wesley Marx. Reports: At the end of the yo-yo — the Federal city. Bulletin of the Atomic Scientists, 21(3):48–50, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1966:FCD

[Mar66a] Howard Margolis. Forum: Civil defense: Notes on Project Harbor. Bulletin of the Atomic Scientists, 22(2):19–21, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1966:RNY

[Mar66b] Howard Margolis. Reports: From New York and Washington: Talking about disarmament. Bulletin of the Atomic Scientists, 22(2):38–40, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1967:UP

[Mar67a] Howard Margolis. The UFO phenomenon. Bulletin of the Atomic Scientists, 23(6):40–42, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sag67, Mun67a].

Martin:1967:BMD

[Mar67b] Laurence W. Martin. Ballistic missile defense and Europe. Bulletin of the Atomic Scientists, 23(5):42–46, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marx:1968:ETP

[Mar68] Wesley Marx. The Eastern Tropical Pacific: Fishing for cooperation. Bulletin of the Atomic Scientists, 24(9):18–22, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mark:1969:MWP

[Mar69] J. Carson Mark. Book review: Men Who Play God, by Norman Moss. Bulletin of the Atomic Scientists, 25(9):44, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshak:1970:RCR

[Mar70] Robert E. Marshak. The Rochester Conferences: The rise of international cooperation in high energy physics. Bulletin of the Atomic Scientists, 26(6):92–98, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1971:CFU

[Mar71a] Charles M. Martin. China: Future of the university. Bulletin of the Atomic Scientists, 27(1):11–15, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marx:1971:FRS

[Mar71b] Wesley Marx. The fall and rise of sewage salvage. Bulletin of the Atomic Scientists, 27(5):10–15, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

March:1972:BBA

[Mar72a] Robert H. March. Books: Public Affairs, by C. P. Snow. Bulletin of the Atomic Scientists, 28(3):45–48, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Margolis:1972:CSE

[Mar72b] George Margolis. Communications: Space emigration. Bulletin of the Atomic Scientists, 28(4):2, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marx:1973:AMM

[Mar73] Wesley Marx. Los Angeles and its mistress machine. Bulletin of the Atomic Scientists, 29(4):4–6, 44–48, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Coo74].

Marcy:1975:FPD

[Mar75] Carl Marcy. Foreign policy debate in a democracy. Bulletin of the Atomic Scientists, 31(9):36–37, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marcus:1976:SWN

[Mar76] Gail H. Marcus. The status of women in the nuclear industry. Bulletin of the Atomic Scientists, 32(4):34–39, April 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Markov:1977:WLT

[Mar77] M. A. Markov. Have we learned to think in a new way? Bulletin of the Atomic Scientists, 33(9):20–23, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bos78].

Marcy:1978:BBW

[Mar78a] Carl Marcy. Books: Nuclear Weapons and World Politics: Alternatives for the Future, by David C. Gompert, Michael Mandelbaum, Richard L. Garwin, John H. Barton. Appendix by F Franklin C. Miller. Bulletin of the Atomic Scientists, 34(2):61– 62, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshak:1978:LOD

[Mar78b] Robert E. Marshak. Letter: Orlov dissident trial in perspective. Bulletin of the Atomic Scientists, 34(7):5–6, September 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1978:LUC

[Mar78c] M. Martin. Letter: Uranium cartel. Bulletin of the Atomic Scientists, 34(2):3–4, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Maw78a].

March:1979:BRB

[Mar79a] Robert March. Book reviews: The Physicists, by Daniel Kevles. Bulletin of the Atomic Scientists, 35(3):61–64, March 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marcy:1979:BRB

[Mar79b] Carl Marcy. Book reviews: The Giants: Russia and America, by Richard J. Barnet. Bulletin of the Atomic Scientists, 35(1): 44, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Markov:1979:RSS

[Mar79c] M. A. Markov. Reflections of a Soviet scientist on Einstein. Bulletin of the Atomic Scientists, 35(3):27–33, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Marcy:1982:EUS

[Mar82a] Carl Marcy. Election 1982: U.S.–Soviet relations. Bulletin of the Atomic Scientists, 38(8):12–14, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marlowe:1982:AQW

[Mar82b] Greg Marlowe. Author query: [W. F. Libby and the history of carbon-14 dating]. Bulletin of the Atomic Scientists, 38(3):44, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1982:IAP

[Mar82c] Gerald E. Marsh. If ‘Atoms for Peace’ are used for war. Bulletin of the Atomic Scientists, 38(2):42, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MM82].

March:1983:F

[Mar83a] Gerald E. March. Furthermore. Bulletin of the Atomic Scientists, 39(6):52–53, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil83, Lei83a].

Mark:1983:WTP

[Mar83b] J. Carson Mark. The weapons tutorial — Part 3: Research, development and production. Bulletin of the Atomic Scientists, 39(3): 45–51, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Parts 1 [Mor83c], 2 [Tsi83i], 4 [Hol83c], 5 [Sar83], and 6 [Rom83].

Marlow:1983:AVD

[Mar83c] Sandra Marlow. Atomic veterans: A daughter’s story. Bulletin of the Atomic Scientists, 39(1):28–29, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1983:FL

[Mar83d] Gerald E. Marsh. Furthermore …. Bulletin of the Atomic Scientists, 39(7):42–43, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GH83].

Marsh:1983:TTB

[Mar83e] Gerald E. Marsh. Threshold Test Ban Treaty: No evidence of cheating. Bulletin of the Atomic Scientists, 39(3):4, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1984:SB

[Mar84] Gerald E. Marsh. Is smaller better? Bulletin of the Atomic Scientists, 40(2):9–10, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1985:BRB

[Mar85a] Gerald Marsh. Book review: Keeping America Uninformed: Government Secrecy in the 1980s, by Donna A. Demac. Bulletin of the Atomic Scientists, 41(4):58, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1985:SSQ

[Mar85b] Gerald E. Marsh. SDI: the stability question. Bulletin of the Atomic Scientists, 41(9):23–24, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1987:PDL

[Mar87a] Gerald E. Marsh. Perspective: Dangers of limited SDI. Bulletin of the Atomic Scientists, 43(2):13–14, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [RC87, SLM87, GM87].

Marshak:1987:PTS

[Mar87b] Robert E. Marshak. Perspective: Time for Sakharov’s global dialogue. Bulletin of the Atomic Scientists, 43(8):7–8, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1988:WP

[Mar88] Gerald E. Marsh. Is war predictable? Bulletin of the Atomic Scientists, 44(8):42, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Markusen:1989:MSC

[Mar89a] Ann R. Markusen. Military spending: Cold War economics. Bulletin of the Atomic Scientists, 45(1):41–44, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gol89].

Marlowe:1989:LMI

[Mar89b] Greg Marlowe. Letter: Mistaken identity. Bulletin of the Atomic Scientists, 45(2):49, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kuz88].

Marples:1989:NSS

[Mar89c] David R. Marples. No soap, say striking Soviet miners. Bulletin of the Atomic Scientists, 45(10):38–40, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marples:1989:RAC

[Mar89d] David R. Marples. Report: After Chernobyl: Ukraine fallout debate. Bulletin of the Atomic Scientists, 45(10):9–11, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1989:BRB

[Mar89e] Gerald E. Marsh. Book review: The Masks of War: American Military Styles in Strategy and Analysis, by Carl H. Builder. Bulletin of the Atomic Scientists, 45(9):46–47, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1989:EUM

[Mar89f] Gerald E. Marsh. From the Editors: U.S. missiles on hair trigger? Bulletin of the Atomic Scientists, 45(4):3, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1990:LCT

[Mar90] David H. Martin. Letter: Canadian tritium to the rescue? No thanks. Bulletin of the Atomic Scientists, 46(2):44–45, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AB89].

Marples:1991:BRE

[Mar91a] David R. Marples. Book review: Eyewitness: The Truth about Chernobyl, by Grigori Medvedev. Bulletin of the Atomic Scientists, 47(5):38–39, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1991:SUD

[Mar91b] Gerald Marsh. START: The ups and downs of downloading. Bulletin of the Atomic Scientists, 47(9):21–23, November 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marples:1993:CLS

[Mar93a] David Marples. Chernobyl’s lengthening shadow. Bulletin of the Atomic Scientists, 49(7):38–43, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BM93].

Marples:1993:LNE

[Mar93b] David R. Marples. Letter: No exaggerated Chernobyl claims from Belarus or Ukraine. Bulletin of the Atomic Scientists, 49(3):47, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [McC93].

Marsh:1994:BRV

[Mar94] Gerald E. Marsh. Book review: The view from space: Submarine Detection from Space: A Study of Russian Capabilities, by Hung P. Nguyen. Bulletin of the Atomic Scientists, 50(5):58, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marples:1996:DDC

[Mar96] David R. Marples. The decade of despair. Bulletin of the Atomic Scientists, 52(3):22–31, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1997:BRD

[Mar97a] Gerald Marsh. Book review: Details, details: The Swords of Armageddon, by Chuck Hansen. Bulletin of the Atomic Scientists, 53(4):54, 56, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

highbeam.com/doc/1G1-19580893.html. See comment [GM87].

Marsh:1997:LMS

[Mar97b] Gerald Marsh. Letter: The myth of the supercomputer. Bulletin of the Atomic Scientists, 53(5):3, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marples:1998:RCU

[Mar98a] David Marples. Report: Chernobyl: Ukraine, 12 years after. Bulletin of the Atomic Scientists, 54(4):15–16, 64, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marsh:1998:RNP

[Mar98b] Gerald Marsh. In response: Nuclear power, yes. Bulletin of the Atomic Scientists, 54(2):47–48, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98].

Marsh:1998:BRT

[Mar98c] Gerald E. Marsh. Book review: Techno-babble: The Future of War: Power, Technology & American World Dominance in the 21st Century, by George Friedman and Meredith Friedman. Bulletin of the Atomic Scientists, 54(1):62–65, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marples:2000:APC

[Mar00] David R. Marples. Atomic-powered Communism. Bulletin of the Atomic Scientists, 56(3):65–66, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/3/65.

Marquardt:2002:MTS

[Mar02a] James J. Marquardt. Not a moment too soon. Bulletin of the Atomic Scientists, 58(1):18, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marquardt:2002:OSM

[Mar02b] James J. Marquardt. Open skies: Not a moment too soon. Bulletin of the Atomic Scientists, 58(1):18–20, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/1/18.

Marsh:2002:TCW

[Mar02c] Gerald Marsh. Twisting Cold war logic. Bulletin of the Atomic Scientists, 58(2):29–31, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/2/29.

Marshall:2006:MK

[Mar06a] Andrew Marshall. Making a killing. Bulletin of the Atomic Scientists, 62(2):36–42, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/2/36.

Marshall:2006:PI

[Mar06b] Andrew Marshall. Plagued by inaction. Bulletin of the Atomic Scientists, 62(4):40–47, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/4/40.

Mason:1948:MPA

[Mas48a] Edward S. Mason. The Marshall Plan and American foreign policy. Bulletin of the Atomic Scientists, 4(1):3–4, 10, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Massey:1948:FDR

[Mas48b] H. S. W. Massey. Freedom demands responsibility — the scientist’s role in society: An English physicist considers his obligations. Bulletin of the Atomic Scientists, 4(11):339–340, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a].

Mason:1957:LEN

[Mas57] Alfred B. Mason. Letter to the Editor: No reckless drivers need apply. Bulletin of the Atomic Scientists, 13(7):273, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Massey:1958:BRS

[Mas58] H. S. W. Massey. British rocket and satellite research. Bulletin of the Atomic Scientists, 14(3):112–114, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Masters:1966:BBS

[Mas66] Roger D. Masters. Books: Deterrence and Strategy, by André Beaufre. Bulletin of the Atomic Scientists, 22(7):38–40, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Masters:1967:CUS

[Mas67] Roger D. Masters. Comments: Uncle Sam, global policeman. Bulletin of the Atomic Scientists, 23(4):32–34, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Masarsky:1976:LNW

[Mas76] Chuck Masarsky. Letter: Nuclear waste power. Bulletin of the Atomic Scientists, 32(4):2, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Masserman:1983:CR

[Mas83] Jules H. Masserman. Commentary: For the record. Bulletin of the Atomic Scientists, 39(2):52, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel82l].

Masserman:1984:CTM

[Mas84] Jules H. Masserman. Commentary: On ‘Three minutes to midnight’. Bulletin of the Atomic Scientists, 40(3):64, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel84e].

Masters:1990:LG

[Mas90] Chris Masters. L’affaire Greenpeace. Bulletin of the Atomic Scientists, 46(2):29, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Masiza:1999:HS

[Mas99a] Zondi Masiza. Horror stories. Bulletin of the Atomic Scientists, 55(3):76–77, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/3/76.

Masiza:1999:ML

[Mas99b] Zondi Masiza. Mandela’s legacy. Bulletin of the Atomic Scientists, 55(2):21–23, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/2/21.

Masco:2008:TA

[Mas08] Joseph Masco. Target audience. Bulletin of the Atomic Scientists, 64(3):22–31, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/64/3/22.

Masco:2013:TN

[Mas13] Joseph P. Masco. Terror as normality. Bulletin of the Atomic Scientists, 69(6):26–32, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/6/26.

Mather:1950:SIS

[Mat50] Kirtley F. Mather. The social implications of science. Bulletin of the Atomic Scientists, 6(7):207–212, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Matson:1956:BBD

[Mat56] Frederick R. Matson. Books: Radiocarbon Dating, by Willard F. Libby. Bulletin of the Atomic Scientists, 12(7):280, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Matthews:1957:CPR

[Mat57] Allan F. Matthews. The cost of preparedness and risk of war. Bulletin of the Atomic Scientists, 13(2):51–55, February 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mattern:1975:LMP

[Mat75] Douglas Mattern. Letters: Misdirected priorities. Bulletin of the Atomic Scientists, 31(6):3–4, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wal74].

Matthias:1979:CWP

[Mat79] Willard C. Matthias. Commentary from Washington: Preview: the Soviet experts and SALT. Bulletin of the Atomic Scientists, 35(4):6–7, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mattern:1980:COW

[Mat80a] Douglas Mattern. Commentary: One world in 1980. Bulletin of the Atomic Scientists, 36(3):60–61, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel80b].

Matthews:1980:CLH

[Mat80b] Allan F. Matthews. Commentary: On ‘learning how to live … ’. Bulletin of the Atomic Scientists, 36(2):63, February 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi80].

Mathieson:1982:NSI

[Mat82a] John A. Mathieson. North–South imbalances. Bulletin of the Atomic Scientists, 38(1):25–28, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mattern:1982:CRS

[Mat82b] Douglas Mattern. Commentary: Requiem for a not so special session. Bulletin of the Atomic Scientists, 38(9):58–59, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathias:1983:TBT

[Mat83a] Charles McC. Mathias, Jr. Test ban treaties: A call for ratification. Bulletin of the Atomic Scientists, 39(5):3–4, May 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathiasen:1983:CCS

[Mat83b] Karl Mathiasen, III. Commentary: Common sense and nuclear peace. Bulletin of the Atomic Scientists, 39(8):65, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZGD83].

Mattern:1984:CPU

[Mat84] Douglas Mattern. Commentary: Put the U.S. back in U.N.E.S.C.O. Bulletin of the Atomic Scientists, 40(5):60–61, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Matsunaga:1985:USS

[Mat85] Spark M. Matsunaga. U.S.–Soviet space cooperation and arms control. Bulletin of the Atomic Scientists, 41(3):17–20, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Matsumure:1988:GE

[Mat88] Takao Matsumure. Ghastly experiments. Bulletin of the Atomic Scientists, 44(8):14, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1997:BEN

[Mat97a] Brendan Mathews. Bulletin: Energy’s NIFty move. Bulletin of the Atomic Scientists, 53(3):10, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1997:BSM

[Mat97b] Brendan Mathews. Bulletin: Surfin’ the .mil domain. Bulletin of the Atomic Scientists, 53(5):9–10, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1997:CCS

[Mat97c] Brendan Mathews. Corona’s cover story. Bulletin of the Atomic Scientists, 53(4):45, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1997:NSI

[Mat97d] Brendan Mathews. No stopping it now. Bulletin of the Atomic Scientists, 53(6):28–31, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Matthews:1997:BCC

[Mat97e] Brendan Matthews. Bulletin: Competitive computing and “creative acquisition”. Bulletin of the Atomic Scientists, 53(3):5–7, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1998:BNS

[Mat98a] Brendan Mathews. Bulletin: Nikitin still in limbo. Bulletin of the Atomic Scientists, 54(2):8, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1998:BSL

[Mat98b] Brendan Mathews. Bulletin: Speaking up or laying low? Bulletin of the Atomic Scientists, 54(6):5–9, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1998:DB

[Mat98c] Brendan Mathews. Deep background. Bulletin of the Atomic Scientists, 54(3):30, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1998:DSH

[Mat98d] Brendan Mathews. Doin’ the Super Hornet dip. Bulletin of the Atomic Scientists, 54(3):28, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mathews:1998:PCJ

[Mat98e] Brendan Mathews. Plane crazy: the Joint Strike Fighter story. Bulletin of the Atomic Scientists, 54(3):26–33, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kel98b].

Mathews:1998:SNN

[Mat98f] Brendan Mathews. Spreading the news on the net. Bulletin of the Atomic Scientists, 54(5):5, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Matthes:2012:EER

[Mat12] Felix Chr. Matthes. Exit economics: The relatively low cost of Germany’s nuclear phase-out. Bulletin of the Atomic Scientists, 68(6):42–54, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/6/42.

Mattick:2018:CAC

[Mat18] Carolyn S. Mattick. Cellular agriculture: The coming revolution in food production. Bulletin of the Atomic Scientists, 74 (1):32–35, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1413059.

Mauck:1950:HCD

[Mau50] Elwyn A. Mauck. History of civil defense in the United States. Bulletin of the Atomic Scientists, 6(8–9):265–270, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Maurer:1977:LRV

[Mau77] Adah Maurer. Letter: Roots of violence. Bulletin of the Atomic Scientists, 33(8):7–8, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav77a].

Mavrodes:1975:LRC

[Mav75] George I. Mavrodes. Letters: Risk and consent. Bulletin of the Atomic Scientists, 31(8):3–4, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys75].

Mawson:1976:LEO

[Maw76] C. A. Mawson. Letter: Eliminating the opposition. Bulletin of the Atomic Scientists, 32(3):4–5, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sal75b, Kee76, Wil76a].

Mawson:1977:LWD

[Maw77] Colin A. Mawson. Letter: Waste disposal options. Bulletin of the Atomic Scientists, 33(2):3–4, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kee76] and comments [Tor77a, Fre77a].

Mawson:1978:LCC

[Maw78a] Colin A. Mawson. Letter: CANDU capacities. Bulletin of the Atomic Scientists, 34(5):4, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar78c].

Mawson:1978:LCF

[Maw78b] Colin A. Mawson. Letter: Capacity figures compared. Bulletin of the Atomic Scientists, 34(3):3, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fes77].

Mayer:1947:LNS

[May47a] J. E. Mayer. Legislation for a National Science Foundation … the bills under consideration. Bulletin of the Atomic Scientists, 3(4–5):124, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mayer:1947:GPW

[May47b] Joseph E. Mayer. Geneva — 1950: A peoples’ world constituent assembly. Bulletin of the Atomic Scientists, 3(12):361–362, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mayer:1952:NSD

[May52] Albert Mayer. The need for synchronized dispersal. Bulletin of the Atomic Scientists, 8(2):49–52, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

May:1985:PVW

[May85] Michael M. May. Perspective: A view from the weapons labs. Bulletin of the Atomic Scientists, 41(6):8–10, June/July 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DeW84].

May:1989:BRB

[May89] Michael M. May. Book review: China Builds the Bomb, by John Lewis and Xue Litai. Bulletin of the Atomic Scientists, 45(3): 42–43, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

May:1990:LIB

[May90] Michael M. May. Letter: If not the bombs, what then? Bulletin of the Atomic Scientists, 46(2):44, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fra89] and comment [Fra90b].

May:2008:TD

[May08] Michael May. The trouble with disarmament. Bulletin of the Atomic Scientists, 64(5):20–21, November/December 2008. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/64/5/20.

May:2017:SFF

[May17] Michael M. May. Safety first: The future of nuclear energy outside the United States. Bulletin of the Atomic Scientists, 73 (1):38–43, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2016.1264210.

Mazrui:1969:AET

[Maz69] Ali A. Mazrui. Africa on the eve of tomorrow: Some social political trends. Bulletin of the Atomic Scientists, 25(9):15–19, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mazur:1981:CSH

[Maz81] Allan Mazur. Commentary: Solar heaters in Israel. Bulletin of the Atomic Scientists, 37(2):56–59, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1948:LPC

[MB48] Byron S. Miller and Harrison S. Brown. Loyalty procedures of the A. E. C. — a report and recommendations. Bulletin of the Atomic Scientists, 42(10):45–48, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miles:1986:LPM

[MB86] Rufus E. Miles, Jr. and Richard K. Betts. Letter: Postwar myth. Bulletin of the Atomic Scientists, 42(7):67–68, August/ September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber86b].

Marin-Bosch:1998:NEN

[MB98] Miguel Marin-Bosch. NATO: Europe’s nuclear family. Bulletin of the Atomic Scientists, 54(1):35–37, January/February 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Macfarlane:1999:CPC

[MB99] Allison Macfarlane and Adam Bernstein. Canning plutonium: Cheaper and faster. Bulletin of the Atomic Scientists, 55(3): 66–69, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/3/66.

Meyerhoff:1962:PE

[MBD^{+}62] Howard A. Meyerhoff, Joseph Berkowitz, DeWitt Stetten, Jr., Robert Gomer, and Paula Fozzy. The publication explosion. Bulletin of the Atomic Scientists, 18(4):33–38, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Woo62].

Manser:1975:LPR

[MC75] W. C. Manser, Jr. and Thomas B. Cochran. Letters: Plutonium recycle. Bulletin of the Atomic Scientists, 31(4):2–3, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [STC74].

McArthur:1965:LEC

[McA65] H. R. McArthur. Letter to the Editor: The Chinese dilemma. Bulletin of the Atomic Scientists, 21(8):29, October 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McAllister:1971:CKE

[McA71] Kenneth McAllister. Communications: Knight errant. Bulletin of the Atomic Scientists, 27(1):52, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McAllister:1976:LPP

[McA76] Kenneth McAllister. Letter: Pathology of power. Bulletin of the Atomic Scientists, 32(7):2, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McArthur:1984:CES

[McA84] H. R. McArthur. Commentary: Euromissiles and short warning time. Bulletin of the Atomic Scientists, 40(10):34–35, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Osg83].

Macfarlane:2007:ONP

[MCAH07] Allison Macfarlane, Michael Corradini, James K. Asselstine, and William G. Halsey. Obstacles to nuclear power. Bulletin of the Atomic Scientists, 63(3):24–25, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/3/24.

McCann:1948:ACD

[McC48] Richard Dyer McCann. Atomic controls: The domestic dilemma. Bulletin of the Atomic Scientists, 4(1):11–13, January 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1951:QC

[MCC51] Brien McMahon, Millard Caldwell, and J. Lawton Collins. Quotes without comment. Bulletin of the Atomic Scientists, 7(7–8):238, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCabe:1954:GR

[McC54a] T. H. McCabe. The Gubser resolutions. Bulletin of the Atomic Scientists, 10(3):71, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McConnell:1954:LEB

[McC54b] R. A. McConnell. Letter to the Editor: The Narrowing Way. Bulletin of the Atomic Scientists, 10(1):33, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McConnell:1954:LES

[McC54c] R. A. McConnell. Letter to the Editor: Spiegel article in April 1954 Bulletin. Bulletin of the Atomic Scientists, 10(6):239, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCleery:1956:RCC

[McC56] Robert S. McCleery. Reader comment: A Christian answer to atomic war. Bulletin of the Atomic Scientists, 12(7):276–277, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Koc57, Dav57a].

McConnell:1957:LEP

[McC57] R. A. McConnell. Letter to the Editor: Proposal to end the testing of (large) bombs. Bulletin of the Atomic Scientists, 13 (1):41, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCorquodale:1965:PS

[McC65] Marjorie K. McCorquodale. Poets and scientists. Bulletin of the Atomic Scientists, 21(9):18–20, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano66j].

McCarter:1969:CLN

[McC69] J. A. McCarter. A Canadian look at nationalization of universities. Bulletin of the Atomic Scientists, 25(5):45–46, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCauley:1981:CNF

[McC81] David McCauley. Commentary: Nuclear freeze in Vermont. Bulletin of the Atomic Scientists, 37(5):64, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McConnaughey:1984:CGR

[McC84] Bayard H. McConnaughey. Commentary: A goodwill race. Bulletin of the Atomic Scientists, 40(6):51, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McClure:1985:LNS

[McC85] Gordon W. McClure. Letter: National security research. Bulletin of the Atomic Scientists, 41(5):55–56, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Das85].

McCally:1986:CESb

[McC86a] Michael McCally. Chernobyl: the emerging story: European worries. Bulletin of the Atomic Scientists, 42(7):47, August/ September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCally:1986:CESa

[McC86b] Michael McCally. Chernobyl: the emerging story: Hospital Number Six: a first-hand report. Bulletin of the Atomic Scientists, 42(7):10–12, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MccGwire:1986:IDD

[Mcc86c] Michael MccGwire. The insidious dogma of deterrence. Bulletin of the Atomic Scientists, 42(10):24–29, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mor87b].

MccGwire:1987:ARS

[Mcc87] Michael MccGwire. Arms race at sea: The changing role of the Soviet Navy. Bulletin of the Atomic Scientists, 43(7):34– 39, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCally:1988:BRB

[McC88a] Michael McCally. Book reviews: Chernobyl: A Russian Journalist’s Eyewitness Account, by Andrey Illesh; Chernobyl: A Novel, by Frederik Pohl. Bulletin of the Atomic Scientists, 44(6):49–50, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCarthy:1988:SSE

[McC88b] Paul McCarthy. Seven steps to expand East–West trade. Bulletin of the Atomic Scientists, 44(5):26–29, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MccGwire:1988:GAP

[Mcc88c] Michael MccGwire. Gorbachev’s arms policy rooted in past. Bulletin of the Atomic Scientists, 44(2):44–46, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCally:1990:RHW

[McC90] Michael McCally. Radiation and health: What the fight is all about. Bulletin of the Atomic Scientists, 46(7):11–14, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano90-54].

McCally:1993:BRE

[McC93] Michael McCally. Book review: Environmental malpractice: Ecocide in the USSR: Health and Nature Under Siege, by Murry Feshbach and Alfred Friendly, Jr. Bulletin of the Atomic Scientists, 49(1):47–48, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mar93b].

McCally:1996:BRA

[McC96] Michael McCally. Book review: Aftereffects: Children of the Atomic Bomb, by James N. Yamazaki; Effects of Atomic Radiation, by William J. Schull. Bulletin of the Atomic Scientists, 52 (2):56–57, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCombie:2009:ESN

[McC09] Charles McCombie. Evaluating solutions to the nuclear waste problem. Bulletin of the Atomic Scientists, 65(6):42–48, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/6/42.

McDonald:1962:SCI

[McD62] Donald McDonald. Science as citizen: [interview with Hans Bethe]. Bulletin of the Atomic Scientists, 18(6):25–28, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McDaniel:1974:AOT

[McD74a] Boyce McDaniel. ‘All in our time’: A physicist at Los Alamos. Bulletin of the Atomic Scientists, 30(10):39–43, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McDonald:1974:IHV

[McD74b] Henry McDonald. Implanting human values into genetic control. Bulletin of the Atomic Scientists, 30(2):21–22, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McDougall:1985:SSR

[McD85] Walter A. McDougall. Sputnik, the space race, and the Cold War. Bulletin of the Atomic Scientists, 41(5):20–25, May 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McDermott:2008:CS

[McD08] Patrice McDermott. The cost of secrecy. Bulletin of the Atomic Scientists, 64(2):64, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/64/2/64.

McDonnell:2013:NPN

[McD13] Timothy McDonnell. Nuclear pursuits: Non-P-5 nuclear-armed states, 2013. Bulletin of the Atomic Scientists, 69(1):62–70, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/1/62.

McEldowney:1952:FOR

[McE52] W. Frank McEldowney. The future of Oak Ridge and Richland: The recommendations of the Scurry Panel on AEC Community Operations. Bulletin of the Atomic Scientists, 8(3):92–93, 96, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McElroy:1970:ARP

[McE70] William D. McElroy. Antarctic research: A pattern of science management. Bulletin of the Atomic Scientists, 26(10):85–88, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

McElroy:1988:BRB

[McE88] Rodney J. McElroy. Book review: Clouds of Secrecy: The Army’s Germ Warfare Tests over Populated Areas, by Leonard A. Cole. Bulletin of the Atomic Scientists, 44(10):46–47, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McGovern:1962:FP

[McG62] George McGovern. Food for peace. Bulletin of the Atomic Scientists, 18(8):44–45, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McGovern:1967:TWA

[McG67] George McGovern. The timid war against hunger. Bulletin of the Atomic Scientists, 23(2):38, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McGuire:1968:CCM

[McG68] Stephen A. McGuire. Current comments: More on the student rebellion. Bulletin of the Atomic Scientists, 24(10):33–34, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab68e].

McGuire:1984:BNB

[McG84] Steven McGuire. Book note: The Forgotten Treaties: A Practical Plan for World Disarmament, by Allan McKnight and Keith Suter. Bulletin of the Atomic Scientists, 40(3):47, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McGraw:1990:RMC

[McG90] Michael McGraw. Report: Military contractors: Another whistleblower down the tubes. Bulletin of the Atomic Scientists, 46(5): 7–8, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McIver:1969:LEP

[McI69] Murdoch McIver. Letter to the Editor: Poetic praise. Bulletin of the Atomic Scientists, 25(9):47, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McKay:1951:CDH

[McK51] H. A. C. McKay. Chemistry [Division at Harwell]. Bulletin of the Atomic Scientists, 7(7–8):230–231, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McKelvsey:1974:BBI

[McK74] V. E. McKelvsey. Books: The International Law of the Ocean Development: Basic Documents, by Shigeru Oda; Ocean Resources and Public Policy, edited by T. Saunders English; The Fate of the Oceans, edited by John J. Logue; Ocean Wealth: Policy and Potential, by George A. Doumani; Mineral Resources of the Deep Seabed. Bulletin of the Atomic Scientists, 30(4):46–48, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McKown:1983:CAV

[McK83] Bryan S. McKown. Commentary: On the atomic veterans. Bulletin of the Atomic Scientists, 39(3):63–64, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kal83b].

McKinney:1986:LTB

[McK86] Thomas A. McKinney. Letter: The test ban and reliability. Bulletin of the Atomic Scientists, 42(6):61–62, June/July 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McKennee:1989:LPS

[McK89] C. T. McKennee. Letter: Provoking Soviets unconstitutional. Bulletin of the Atomic Scientists, 45(4):52–53, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [The89f].

McKiernan:1999:TWK

[McK99] Kevin McKiernan. Turkey’s war on the Kurds. Bulletin of the Atomic Scientists, 55(2):26–37, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/2/26.

McKibben:2007:TPO

[McK07] Bill McKibben. Thinking past ourselves. Bulletin of the Atomic Scientists, 63(6):28–31, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/6/28.

McKenzie:2022:HBM

[McK22] Jessica McKenzie. How bitcoin makes burning fossil fuels more profitable than ever. Bulletin of the Atomic Scientists, 78(4): 203–207, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2087379.

McKenzie:2023:BME

[McK23a] Jessica McKenzie. Bill McKibben explains what individuals can do to win the climate fight. Together. Bulletin of the Atomic Scientists, 80(1):9–13, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2293566.

McKenzie:2023:WAE

[McK23b] Jessica McKenzie. “The world has already ended”: Britt Wray on living with the horror and trauma of climate crisis. Bulletin of the Atomic Scientists, 80(1):25–31, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2293571.

McKenzie:2025:FIT

[McK25a] Jessica McKenzie. ‘Fragile, impermanent things’: Joseph Tainter on what makes civilizations fall. Bulletin of the Atomic Scientists, 81(2):114–120, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2025.2464444.

McKenzie:2025:IAE

[McK25b] Jessica McKenzie. Introduction: (almost) everything you wanted to know about tipping points, but were too afraid to ask. Bulletin of the Atomic Scientists, 81(2):81–82, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2025.2465146.

McLean:1949:BBP

[McL49] Franklin McLean. Books: Peace or Pestilence, by Theodore Rosebury. Bulletin of the Atomic Scientists, 5(12):353–354, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McLean:1953:PER

[McL53] Joseph E. McLean. Project East River — survival in the Atomic Age. Bulletin of the Atomic Scientists, 9(7):247–252, 288, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McLean:1965:LFK

[McL65] Louis A. McLean. Letters: Fish kill. Bulletin of the Atomic Scientists, 21(2):35, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McLaren:1966:SCT

[McL66] Robert Bruce McLaren. Science and contemporary theology. Bulletin of the Atomic Scientists, 22(3):25–26, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McLeish:2011:SQE

[McL11] Caitríona McLeish. Status quo or evolution: What next for the intersessional process of the biological and toxin weapons convention? Bulletin of the Atomic Scientists, 67(3):34–43, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/3/34.

McLeman:2018:MDR

[McL18] Robert McLeman. Migration and displacement risks due to mean sea-level rise. Bulletin of the Atomic Scientists, 74(3): 148–154, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1461951.

McMahon:1948:AAE

[McM48a] Brien McMahon. Author of Atomic Energy Act comments on Governor Dewey’s speech. Bulletin of the Atomic Scientists, 4 (10):298–299, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1948:OIC

[McM48b] Brien McMahon. Our international control plan not sufficiently known to the world. Bulletin of the Atomic Scientists, 4(9):262, 288, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1949:CBB

[McM49a] Brien McMahon. Comment on Blackett’s book. Bulletin of the Atomic Scientists, 5(2):40–42, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1949:SWR

[McM49b] Brien McMahon. Should we reveal the size of our atomic stockpile? Bulletin of the Atomic Scientists, 5(3):66–68, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1950:NBA

[McM50] Brian McMahon. A new bid for atomic peace. Bulletin of the Atomic Scientists, 6(3):80–82, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1951:AWD

[McM51] Brien McMahon. Atomic weapons and defense. Bulletin of the Atomic Scientists, 7(10):297–301, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1952:SRI

[McM52] Brien McMahon. Survival — the real issue of our times. Bulletin of the Atomic Scientists, 8(6):173–175, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahon:1954:AAA

[McM54] Brien McMahon. [An atomic Army and an atomic Navy and an atomic Air Force]. Bulletin of the Atomic Scientists, 10(8): 317, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMurphy:1958:LEU

[McM58] William A. McMurphy. Letter to the Editor: Problems of peace. Bulletin of the Atomic Scientists, 14(4):156, 161, April 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMahan:1965:RBM

[McM65] Richard H. McMahan, Jr. Rationales for ballistic missile defense policy. Bulletin of the Atomic Scientists, 21(3):37–40, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys64, Dys65].

McMahon:1967:IAE

[McM67] Brian McMahon. International atomic energy: 1947. Bulletin of the Atomic Scientists, 23(3):46, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Short excerpt from [Sen47].

McMahan:1982:CNM

[McM82] Jefferson McMahan. Commentary: On nuclear modernization in Europe. Bulletin of the Atomic Scientists, 38(8):57, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMillan:1994:SFF

[McM94a] Priscilla Johnson McMillan. The Sudoplatov file: Flimsy memories. Bulletin of the Atomic Scientists, 50(4):30–33, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-15587910.
html.

McMillan:1994:TWF

[McM94b] Priscilla Johnson McMillan. They weren’t “friends”. Bulletin of the Atomic Scientists, 50(4):31, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McMillan:2003:ETO

[McM03] Priscilla Johnson McMillan. Edward Teller: Obituary. Bulletin of the Atomic Scientists, 59(6):10–11, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-110621486.
html.

McMillan:2005:RJR

[McM05] Priscilla Johnson McMillan. The ruin of J. Robert Oppenheimer, and the birth of the modern arms race. Viking, New York, NY, USA, 2005. ISBN 0-670-03422-3. viii + 373 + 16 pp. LCCN QC16.O62 M36 2005.

McNaughton:1948:CFS

[McN48] A. G. L. McNaughton. Canada favors suspension of UN AEC. Bulletin of the Atomic Scientists, 4(7):221, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNear:1959:STA

[McN59] Everett McNear. Some thoughts about the painter’s craft. Bulletin of the Atomic Scientists, 15(2):60–63, February 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNeil:1962:IU

[McN62] Elton B. McNeil. An international university. Bulletin of the Atomic Scientists, 18(8):23–24, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNamara:1963:TBM

[McN63a] Robert S. McNamara. Test ban: The military argument. Bulletin of the Atomic Scientists, 19(8):42–43, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNeil:1963:CWR

[McN63b] Elton B. McNeil. The children’s war revisited. Bulletin of the Atomic Scientists, 19(2):22–24, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNeil:1963:FIR

[McN63c] Elton B. McNeil. Forum: International relations and the psychologist: The scientific gulf. Bulletin of the Atomic Scientists, 19(9):33–35, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNamara:1964:NSN

[McN64] Robert S. McNamara. National security and nuclear strength. Bulletin of the Atomic Scientists, 20(3):35–39, March 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNamara:1967:RSD

[McN67a] Robert S. McNamara. Remarks by the Secretary of Defense, September 18, 1967. Bulletin of the Atomic Scientists, 23(10):26– 31, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNamara:1967:USD

[McN67b] Robert S. McNamara. U.S. Secretary of Defense testifies. Bulletin of the Atomic Scientists, 23(6):21–24, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNamara:1972:CT

[McN72] Robert S. McNamara. A critical truth. Bulletin of the Atomic Scientists, 28(7):39–43, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNary:1976:LCP

[McN76] Thomas A. McNary. Letter: Clear but perturbing. Bulletin of the Atomic Scientists, 32(10):6, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lif76].

McNamara:1977:PIA

[McN77] Robert S. McNamara. A pittance for international aid. Bulletin of the Atomic Scientists, 33(9):36–38, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNeil:1982:PPH

[McN82] William H. McNeil. The pursuit of power: a historian reflects. Bulletin of the Atomic Scientists, 38(4):19–23, April 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNeil:2009:NS

[McN09] Scott E. McNeil. Nanomaterial safety. Bulletin of the Atomic Scientists, 65(1):56–61, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/65/1/56.

McPherson:1965:SFT

[McP65] Archibald T. McPherson. Synthetic food for tomorrow’s billions. Bulletin of the Atomic Scientists, 21(7):6–11, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McReynolds:1951:NCE

[McR51] John W. McReynolds. Noted on the compleat expert. Bulletin of the Atomic Scientists, 7(1):15–16, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ley51]. Reprinted from The American Scientist, April 1950.

McVay:1971:DWM

[McV71] Scott McVay. Does the whale’s magnitude diminish? — will he perish? Bulletin of the Atomic Scientists, 27(2):38–41, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mattison:1964:NFP

[MD64] Lindsay Mattison and Richard Daly. Nevada fallout: Past and present: Hazards. Bulletin of the Atomic Scientists, 20(4):41–45, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and response [DMD64].

Moravcsik:1966:LEI

[MD66] Michael J. Moravcsik and Erick Dreikurs. Letter to the Editor: On improving communication. Bulletin of the Atomic Scientists, 22(5):31, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maddox:1975:LLS

[MD75] John Maddox and Samuel H. Day, Jr. Letters: The lost sheep. Bulletin of the Atomic Scientists, 31(7):4, September 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75f, SG75].

Moore:1981:CAN

[MD81] Robert Moore and Ann Morrissett Davidon. Commentary: On ‘Anti-nuclear connections’. Bulletin of the Atomic Scientists, 37 (8):61–62, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nel81].

Martinez-Diaz:2018:IRT

[MD18] Leonardo Martinez-Diaz. Investing in resilience today to prepare for tomorrow’s climate change. Bulletin of the Atomic Scientists, 74(2):66–72, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2018.1436805.

Mecklin:2024:ISE

[MD24] John Mecklin and Dan Drollette, Jr. Introduction: Securing elections, democracy, and the information ecosystem in a critical political year. Bulletin of the Atomic Scientists, 80(5): 273–274, 2024. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2024.2395112.

Markusen:1981:REH

[MDB81] Eric Markusen, Jeffrey Dunham, and Ronald Bee. 1980 Rabinowitch Essay on how to eliminate the threat of nuclear war: A nuclear education plan. Bulletin of the Atomic Scientists, 37(5): 39–42, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Merriam:1951:CHS

[ME51] Robert E. Merriam and Norman Elkin. Cities are here to stay. Bulletin of the Atomic Scientists, 7(9):251–255, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McNaughton:1947:AFR

[MEA^{+}47] General Andrew G. L. McNaughton, Jacques Errera, Captain Alvara Alberto, Bertrand Goldschmidt, and John E. Vance. Agency’s functions in relation to location and mining of ores. Bulletin of the Atomic Scientists, 3(9):257–258, September 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meadow:1962:BBF

[Mea62] Kathryn P. Meadow. Books: The Future of Our Cities, by Robert A. Futterman. Bulletin of the Atomic Scientists, 18(1):41, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mead:1969:PPB

[Mea69] Margaret Mead. Public policy and behavioral science. Bulletin of the Atomic Scientists, 25(10):8–10, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mead:1972:LYE

[Mea72] Margaret Mead. A loved yet endangered planet. Bulletin of the Atomic Scientists, 28(7):46–, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mearsheimer:1985:PCD

[Mea85] John J. Mearsheimer. Prospects for conventional deterrence in Europe. Bulletin of the Atomic Scientists, 41(7):158–162, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lam85].

Mearsheimer:1993:BRM

[Mea93] John J. Mearsheimer. Book review: McNamara’s war: Promise and Power: The Life and Times of Robert McNamara, by Deborah Shapley . Bulletin of the Atomic Scientists, 49(6):49–51, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kap93].

Meckling:1963:EST

[Mec63] William H. Meckling. Economics and space technology. Bulletin of the Atomic Scientists, 19(5):15–17, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2012:GNE

[Mec12] John Mecklin. The German nuclear exit: Introduction. Bulletin of the Atomic Scientists, 68(6):6–9, November/December 2012. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/68/6/6.

Mecklin:2013:IFN

[Mec13a] John Mecklin. Introduction: a French nuclear exit? Bulletin of the Atomic Scientists, 69(1):9–10, January/February 2013. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/69/1/9.

Mecklin:2013:IUN

[Mec13b] John Mecklin. Introduction: US nuclear exit? Bulletin of the Atomic Scientists, 69(2):9–11, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/2/9.

Mecklin:2014:CHW

[Mec14] John Mecklin. California here we come? Bulletin of the Atomic Scientists, 70(5):24–25, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2015:IBY

[Mec15a] John Mecklin. Introduction: The Bulletin at the young age of 70. Bulletin of the Atomic Scientists, 71(1):10–12, January 2015. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2015:BBL

[Mec15b] John Mecklin. On its 70th birthday, the Bulletin looks forward. Bulletin of the Atomic Scientists, 71(6):1–3, November 2015. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2016:EN

[Mec16a] John Mecklin. Editor’s note. Bulletin of the Atomic Scientists, 72 (2):116, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2016:FCC

[Mec16b] John Mecklin. Former covert CIA operations officer Valerie Plame. Bulletin of the Atomic Scientists, 72(2):74–78, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2016:FUA

[Mec16c] John Mecklin. Former US ambassador to Russia Michael McFaul. Bulletin of the Atomic Scientists, 72(3):135–139, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2016:IIS

[Mec16d] John Mecklin. Introduction: International security in the age of renewables. Bulletin of the Atomic Scientists, 72(6):377, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2016:IPN

[Mec16e] John Mecklin. Introduction: Practical nuclear questions for the candidates in an unusual Presidential election. Bulletin of the Atomic Scientists, 72(5):273–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2016:UAN

[Mec16f] John Mecklin. UAE Ambassador and nuclear power expert Hamad Alkaabi. Bulletin of the Atomic Scientists, 72(1):25–31, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mecklin:2017:IA

[Mec17a] John Mecklin. Introduction: Into the aftermath. Bulletin of the Atomic Scientists, 73(4):210, 2017. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2017.
1338000.

Mecklin:2017:INP

[Mec17b] John Mecklin. Introduction: Nuclear power and the urgent threat of climate change. Bulletin of the Atomic Scientists, 73(1):1, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2016.1265351.

Mecklin:2017:IET

[Mec17c] John Mecklin. Introduction: The evolving threat of hybrid war. Bulletin of the Atomic Scientists, 73(5):298–299, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1362905.

Mecklin:2017:ITA

[Mec17d] John Mecklin. Introduction: The Trump administration’s approach to existential threat. Bulletin of the Atomic Scientists, 73 (6):368–369, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1388663.

Mecklin:2018:IGN

[Mec18a] John Mecklin. Introduction: Good news in perilous times. Bulletin of the Atomic Scientists, 74(1):1, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2017.
1413053.

Mecklin:2018:IGM

[Mec18b] John Mecklin. Introduction: The great missile defense dilemma. Bulletin of the Atomic Scientists, 74(4):197–198, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1486571.

Mecklin:2018:INT

[Mec18c] John Mecklin. Introduction: The new threat matrix. Bulletin of the Atomic Scientists, 74(6):367, 2018. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2018.
1533164.

Mecklin:2019:CWR

[Mec19a] John Mecklin. Christie Whitman on Republicans and climate change. Bulletin of the Atomic Scientists, 75(2):63–65, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1580875.

Mecklin:2019:DRA

[Mec19b] John Mecklin. Dealing realistically with the artificial intelligence revolution. Bulletin of the Atomic Scientists, 75(3):93– 94, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1604812.

Mecklin:2019:FDS

[Mec19c] John Mecklin. Former Defense Secretary William Perry: Why we must describe doomsday to keep it from happening. Bulletin of the Atomic Scientists, 75(6):265–268, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1680046.

Mecklin:2019:ICC

[Mec19d] John Mecklin. Introduction: Climate change action — from the right. Bulletin of the Atomic Scientists, 75(2):57–58, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1580867.

Mecklin:2019:IWD

[Mec19e] John Mecklin. Introduction: The wasteful and dangerous worldwide nuclear modernization craze. Bulletin of the Atomic Scientists, 75(1):1–2, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1555973.

Mecklin:2019:JWL

[Mec19f] John Mecklin. Jon Wolfsthal on the link between nuclear strategy and the nuclear modernization budget. Bulletin of the Atomic Scientists, 75(1):9–13, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1555990.

Mecklin:2019:STT

[Mec19g] John Mecklin. Stories of technological threat — and hope. Bulletin of the Atomic Scientists, 75(6):263–264, 2019. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1680045.

Mecklin:2019:WSW

[Mec19h] John Mecklin. Why Star Wars should remain a cinematic fantasy. Bulletin of the Atomic Scientists, 75(4):135–136, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1630101.

Mecklin:2020:BFH

[Mec20a] John Mecklin. Beatrice Fihn: How to implement the nuclear weapons ban treaty. Bulletin of the Atomic Scientists, 76(6): 336–340, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1846423.

Mecklin:2020:CNN

[Mec20b] John Mecklin. Can the nuclear nonproliferation regime be saved when arms control is collapsing? Bulletin of the Atomic Scientists, 76(2):61–62, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1730605.

Mecklin:2020:CCA

[Mec20c] John Mecklin. Climate change action requires … actual action. Bulletin of the Atomic Scientists, 76(5):227, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1811485.

Mecklin:2020:CCE

[Mec20d] John Mecklin. The climate change evidence right before our eyes. And a note on COVID-19. Bulletin of the Atomic Scientists, 76(3):119–120, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1753318.

Mecklin:2020:IDF

[Mec20e] John Mecklin. An innovative and determined future for the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 76(6):277–279, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1847442.

Mecklin:2020:NCL

[Mec20f] John Mecklin. Nobel chemistry laureate Jennifer Doudna on the promise and peril of the genetic editing revolution. Bulletin of the Atomic Scientists, 76(6):285–289, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1846407.

Mecklin:2020:SHR

[Mec20g] John Mecklin. Siegfried Hecker on remembering history while planning the future of nuclear arms control. Bulletin of the Atomic Scientists, 76(6):309–314, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1846415.

Mecklin:2020:WNW

[Mec20h] John Mecklin. Why nuclear weapons should be a major focus of the 2020 campaign. Bulletin of the Atomic Scientists, 76(1):1–2, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1702274.

Mecklin:2021:AMH

[Mec21a] John Mecklin. Alan Miller: How the News Literacy Project teaches schoolchildren (and adults) to dismiss and debunk Internet disinformation. Bulletin of the Atomic Scientists, 77(3): 111–115, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1912002.

Mecklin:2021:HRR

[Mec21b] John Mecklin. How the renewables revolution can move from catchphrase to reality. Bulletin of the Atomic Scientists, 77 (6):279–280, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1992869.

Mecklin:2021:IDR

[Mec21c] John Mecklin. Interview: Diane Randall, Director General of the Friends Committee on National Legislation, discusses restraining the US defense budget. Bulletin of the Atomic Scientists, 77(5):244–248, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1964233.

Mecklin:2021:ITC

[Mec21d] John Mecklin. Interview: Tom Collina of the Ploughshares Fund on the politics of defense spending. Bulletin of the Atomic Scientists, 77(5):255–258, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1964250.

Mecklin:2021:IAN

[Mec21e] John Mecklin. Introduction: Advice for a new administration facing difficult times. Bulletin of the Atomic Scientists, 77(1):1–2, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1862514.

Mecklin:2021:ICW

[Mec21f] John Mecklin. Introduction: Can we make overspending on the military politically costly? Bulletin of the Atomic Scientists, 77(5):237–238, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1964219.

Mecklin:2021:ICA

[Mec21g] John Mecklin. Introduction: Climate action in the general interest. Bulletin of the Atomic Scientists, 77(2):65–66, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1888466.

Mecklin:2021:IHD

[Mec21h] John Mecklin. Introduction: How to dial back a disinformation dystopia. Bulletin of the Atomic Scientists, 77(3):109–110, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1914929.

Mecklin:2022:IRL

[Mec22a] John Mecklin. Interview: Robert Latiff on the worsening international security situation in space. Bulletin of the Atomic Scientists, 78(1):3–5, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.2014228.

Mecklin:2022:IGA

[Mec22b] John Mecklin. Interview with Graham Allison: Are the United States and China charging into Thucydides’s trap? Bulletin of the Atomic Scientists, 78(2):53–58, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2041839.

Mecklin:2022:ICU

[Mec22c] John Mecklin. Introduction: Can the United States and China co-exist in the 21st century? Will they? Bulletin of the Atomic Scientists, 78(2):51–52, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2022.2041835.

Mecklin:2022:WFF

[Mec22d] John Mecklin. Why the final frontier should not become the final battleground. Bulletin of the Atomic Scientists, 78(1):1–2, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.2020988.

Mecklin:2023:EIC

[Mec23a] John Mecklin. An extended interview with Christopher Nolan, director of Oppenheimer. Bulletin of the Atomic Scientists, 79 (4):203–209, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2223070.

Mecklin:2023:ICC

[Mec23b] John Mecklin. Interview: California Congressman Ted Lieu on what you, as a citizen, can do about existential threats.

Bulletin of the Atomic Scientists, 80(1):14–16, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2293567.

Mecklin:2023:IEM

[Mec23c] John Mecklin. Interview: Emerging military technology expert Paul Scharre on global power dynamics in the AI age. Bulletin of the Atomic Scientists, 79(5):284–288, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2245240.

editor-in-chief:2023:IHP

[Mec23d] John Mecklin. Introduction: The hype, peril, and promise of artificial intelligence. Bulletin of the Atomic Scientists, 79(5): 283, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00963402.2023.2246264.

Mecklin:2024:IAE

[Mec24] John Mecklin. An interview about the 2024 election with Harper Reed, chief technology officer for Obama 2012. Bulletin of the Atomic Scientists, 80(5):297–301, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2388465.

MBR:1947:MBR

[Med47] Medical Board of Review. Medical Board of Review reports to U. S. A. E. C. Bulletin of the Atomic Scientists, 3(9):272–274, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Medwed:1976:TTE

[Med76] Howard Medwed. Third thoughts on Entebbe. Bulletin of the Atomic Scientists, 32(9):8–9, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Medawar:1981:CAY

[Med81] Sir Peter B. Medawar. Commentary: Advice to a young scientist. Bulletin of the Atomic Scientists, 37(4):51–54, April 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Medler:1991:PB

[Med91] Alex Medler. Paying the bill. Bulletin of the Atomic Scientists, 47 (8):38, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mehic:1995:WDY

[Meh95] Dzenita Mehic. “We are dying of your protection”. Bulletin of the Atomic Scientists, 51(2):41–44, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kri95].

Mehta:2021:HBC

[Meh21] Rupal Mehta. How Biden can say goodbye to “America First” on nuclear issues. Bulletin of the Atomic Scientists, 77(1):11– 12, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1859859.

Meier:1948:WSA

[Mei48] Richard L. Meier. What should the atomic scientists do now? Bulletin of the Atomic Scientists, 4(3):81–82, March 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1951:LEB

[Mei51a] R. L. Meier. Letter to the Editor: Long-Term Prospects for Essential Minerals. Bulletin of the Atomic Scientists, 7(10):312, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1951:HML

[Mei51b] Richard L. Meier. How much is left in the hat? Bulletin of the Atomic Scientists, 7(7–8):205, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1951:LTP

[Mei51c] Richard L. Meier. The long-term prospects for essential minerals. Bulletin of the Atomic Scientists, 7(7–8):214–216, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mei51a] for a table that was lost during editing of this article.

Meier:1951:OSS

[Mei51d] Richard L. Meier. The origins of the scientific species. Bulletin of the Atomic Scientists, 7(6):169–173, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1952:BRB

[Mei52] R. L. Meier. Book review: World Population and Future Resources. Bulletin of the Atomic Scientists, 8(5):163–164, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meiklejohn:1953:IUH

[Mei53] Alexander Meiklejohn. Integrity of the universities — how to defend it. Bulletin of the Atomic Scientists, 9(5):193–194, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1954:AED

[Mei54] Richard L. Meier. Automatic and economic development. Bulletin of the Atomic Scientists, 10(4):129–133, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1956:BAS

[Mei56] R. L. Meier. Beyond atomic stalemate. Bulletin of the Atomic Scientists, 12(5):147–152, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1961:LER

[Mei61] Richard L. Meier. Letter to the Editor: Is reactor development economical? Bulletin of the Atomic Scientists, 17(2):40, 81, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WW60] and response [WW61].

Meitner:1964:LB

[Mei64] Lise Meitner. Looking back. Bulletin of the Atomic Scientists, 20(9):2–7, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1968:BBY

[Mei68] R. L. Meier. Books: The Year 2000, by Herman Kahn and Anthony J. Wiener. Bulletin of the Atomic Scientists, 24(1):51–52, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1969:SIN

[Mei69] Richard L. Meier. The social impact of a nuplex. Bulletin of the Atomic Scientists, 25(3):16–21, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1970:MTR

[Mei70] Richard L. Meier. The metropolis and the transformation of resources. Bulletin of the Atomic Scientists, 26(5):2–5, 36–37, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meisel:1973:VCS

[Mei73] Albert Meisel. A voice of conscience is stilled. I. Scientist and humanist. Bulletin of the Atomic Scientists, 29(6):4, June 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1977:BWH

[Mei77a] Richard L. Meier. Books: Whither human settlements? The Home of Man, by Barbara Ward. Bulletin of the Atomic Scientists, 33(5):62–64, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:1977:MAS

[Mei77b] Richard L. Meier. Multinationals as agents of social development. Bulletin of the Atomic Scientists, 33(9):30–32, 34–35, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sti78].

Meier:1990:BRS

[Mei90] Paul Meier. Book review: Sly statistics: Deadly Deceit: Low-Level Radiation, High-Level Cover-up, by Jay M. Gould and Benjamin A. Goldman. Bulletin of the Atomic Scientists, 46(7):40– 41, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [GM90].

Meir:1998:LIS

[Mei98] Shemuel Meir. Letter: Israel should ratify the chemical treaty. Bulletin of the Atomic Scientists, 54(4):5, 65, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meier:2001:BTN

[Mei01] Oliver Meier. Bioweapons treaty: Neither trust nor verify, says U.S. Bulletin of the Atomic Scientists, 57(6):19–21, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/6/19.

Meier:2006:NTN

[Mei06] Oliver Meier. NATO: Tied in nuclear knots. Bulletin of the Atomic Scientists, 62(1):14–16, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/14.

Meier:2020:WGW

[Mei20] Oliver Meier. Why Germany won’t build its own nuclear weapons and remains skeptical of a Eurodeterrent. Bulletin of the Atomic Scientists, 76(2):76–84, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1728967.

Meltzer:1953:IFAb

[Mel53a] Bernard Meltzer. Invoking the Fifth Amendment — a rejoinder. Bulletin of the Atomic Scientists, 9(5):185–186, 200, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meltzer:1953:IFAa

[Mel53b] Bernard D. Meltzer. Invoking the Fifth Amendment — some legal and practical considerations. Bulletin of the Atomic Scientists, 9(5):176–180, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Melman:1958:HCI

[Mel58] Seymour Melman. How can inspection be made to work? Bulletin of the Atomic Scientists, 14(7):270–272, September 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Melman:1971:AMI

[Mel71] Seymour Melman. After the military–industrial complex? Bulletin of the Atomic Scientists, 27(3):7–9, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Melko:1975:PSW

[Mel75] Matthew Melko. Peace: A subject worth studying: It may be the more normal human condition. Bulletin of the Atomic Scientists, 31(4):31–34, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and response [LM75, Gla75d].

Mellman:1978:CHR

[Mel78] Mark Mellman. Commentary: Human rights: a different perspective. Bulletin of the Atomic Scientists, 34(9):60–61, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zim78] and comment [Zim79].

Mello:1997:NBN

[Mel97] Greg Mello. New bomb, no mission. Bulletin of the Atomic Scientists, 53(3):28–32, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mello:2000:ODF

[Mel00] Greg Mello. That old designing fever. Bulletin of the Atomic Scientists, 56(1):51–57, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/51.

MCPL:1979:CSF

[Mem79] Members of Congress for Peace through Law. Commentary: Strike force for the Middle East? Bulletin of the Atomic Scientists, 35(8):64, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

MCPL:1980:CPD

[Mem80] Members of Congress for Peace through Law. Commentary: On Presidential Directive #59. Bulletin of the Atomic Scientists, 36 (9):63–64, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Menke:1948:NFS

[Men48] John R. Menke. Nuclear fission as a source of power. Bulletin of the Atomic Scientists, 4(4):115–120, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Menke:1953:WDP

[Men53] John R. Menke. Why the dual purpose breeder reactor? Bulletin of the Atomic Scientists, 9(5):162–164, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendlovitz:1964:TWP

[Men64] Saul H. Mendlovitz. Teaching war prevention. Bulletin of the Atomic Scientists, 20(2):19–22, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mencher:1968:MGS

[Men68] Alan G. Mencher. Management by government: Science and technology in Britain. Bulletin of the Atomic Scientists, 24(5):22–27, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mencher:1969:RSA

[Men69] Alan G. Mencher. Reports: Scientists among diplomats. Bulletin of the Atomic Scientists, 25(1):46–48, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mencher:1971:ISD

[Men71] Alan G. Mencher. IV. On the social deployment of science. Bulletin of the Atomic Scientists, 27(10):34–38, December 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1986:ACB

[Men86] Jack Mendelsohn. Arms control ball in Reagan’s court. Bulletin of the Atomic Scientists, 42(10):8–10, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1989:SDC

[Men89] Jack Mendelsohn. START deals cut at Jackson Hole. Bulletin of the Atomic Scientists, 45(10):25–27, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1991:NWO

[Men91a] Jack Mendelsohn. New world order: Rhetoric v. reality. Bulletin of the Atomic Scientists, 47(5):34–35, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1991:RSG

[Men91b] Jack Mendelsohn. Report: START at a glance. Bulletin of the Atomic Scientists, 47(1):9, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1991:RSS

[Men91c] Jack Mendelsohn. Report: START: Senate will grouse, then ratify. Bulletin of the Atomic Scientists, 47(1):8–10, January/ February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendlovitz:1991:BRM

[Men91d] Saul Mendlovitz. Book review: Moynihan’s world: On the Law of Nations, by Daniel Patrick Moynihan. Bulletin of the Atomic Scientists, 47(9):42–43, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1992:EFS

[Men92a] Jack Mendelsohn. Enroll in five step program. Bulletin of the Atomic Scientists, 48(4):40–41, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelson:1992:BRD

[Men92b] Sarah E. Mendelson. Book review: Diplomat games: Untying the Afghan Knot: Negotiating Soviet Withdrawal, by Riaz M. Khan. Bulletin of the Atomic Scientists, 48(9):43–44, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1993:HHH

[Men93a] Jack Mendelsohn. Huddling with the honchos in Havana. Bulletin of the Atomic Scientists, 49(7):14–19, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1993:WAR

[Men93b] Jack Mendelsohn. The world according to Raúl [Castro]. Bulletin of the Atomic Scientists, 49(7):16, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1995:GOS

[Men95] Jack Mendelsohn. Guest opinion: Stiff-arming Russia. Bulletin of the Atomic Scientists, 51(4):4, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1996:D

[Men96a] Jack Mendelsohn. “Demarcation”. Bulletin of the Atomic Scientists, 52(1):29, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1996:TIS

[Men96b] Jack Mendelsohn. A tenth inning for Star Wars: The ABM Treaty is in DANGER, says an arms control expert. Bulletin of the Atomic Scientists, 52(1):24, 28–31, January/February 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Menon:1997:OFS

[Men97] Rajan Menon. The once and future superpower. Bulletin of the Atomic Scientists, 53(1):29–34, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1999:MDI

[Men99] Jack Mendelsohn. Missile defense: and it still won’t work. Bulletin of the Atomic Scientists, 55(3):29–33, May/June 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/3/29.

Mendelsohn:2000:SB

[Men00] Jack Mendelsohn. Still bound. Bulletin of the Atomic Scientists, 56(1):42–43, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/42.

Merck:1946:ORB

[Mer46a] George W. Merck. Official report on biological warfare. Bulletin of the Atomic Scientists, 2(7–8):16–18, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Merriam:1946:PP

[Mer46b] Charles E. Merriam. Physics and politics. Bulletin of the Atomic Scientists, 1(10):9–11, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Merton:1947:RRS

[Mer47] Robert K. Merton. Replies to Ridenour: The seven propositions of Professor Ridenour. Bulletin of the Atomic Scientists, 3(8): 225, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mercer:1954:LEM

[Mer54a] Mary L. Mercer. Letter to the Editor: Madame Joliot and the American Chemical Society. Bulletin of the Atomic Scientists, 10 (6):213, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Merrill:1954:LEP

[Mer54b] Herbert M. Merrill. Letter to the Editor: Panic, psychology, and the bomb. Bulletin of the Atomic Scientists, 10(4):145, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wyl54, Mar54c].

Merriam:1972:CSE

[Mer72] Marshal F. Merriam. Communications: Solar energy. Bulletin of the Atomic Scientists, 28(10):2–3, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alf72].

Merriam:1979:BRB

[Mer79] Marshal F. Merriam. Book review: Wind Power, and Other Energy Sources, by David Rittenhouse Inglis. Bulletin of the Atomic Scientists, 35(9):52, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Merrill:1990:BRF

[Mer90] Dennis Merrill. Book review: Flexing U.S. muscle: Confronting the Third World: United States Foreign Policy 1945–1980, by Gabriel Kolko. Bulletin of the Atomic Scientists, 46(2):42–43, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meredith:2024:WGC

[Mer24] Michael Meredith. When glaciers calve: Large underwater tsunamis discovered at edge of Antarctica, likely affecting ice melt, climate and marine ecosystem. Bulletin of the Atomic Scientists, 80(4):235–238, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2364506.

Meselson:1964:BBW

[Mes64a] Matthew Meselson. Books: Tomorrow’s Weapons, Chemical and Biological, by Jacquard Hirshorn Rothschild. Bulletin of the Atomic Scientists, 20(8):35–36, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mesthene:1964:BBN

[Mes64b] Emmanuel G. Mesthene. Books: Scientists and National Policy-Making, edited by Robert Gilpin and Christopher Wright. Bulletin of the Atomic Scientists, 20(9):25–27, November 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meselson:1970:BNP

[Mes70] Matthew S. Meselson. Behind the Nixon policy for chemical and biological warfare: Background report. Bulletin of the Atomic Scientists, 26(1):23–24, 26–34, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meselson:1971:TGV

[Mes71] Matthew Meselson. Tear gas in Vietnam and the return of poison gas. Bulletin of the Atomic Scientists, 27(3):17–19, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meselson:1972:GWG

[Mes72] Matthew Meselson. Gas warfare and the Geneva Protocol of 1925. Bulletin of the Atomic Scientists, 28(2):33–37, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mesic:1982:BRB

[Mes82] Penelope Mesic. Book review: Riddley Walker, by Russell Hoban. Bulletin of the Atomic Scientists, 38(6):49–50, June 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mesic:1983:BRB

[Mes83a] Penelope Mesic. Book reviews: Hiroshima No Pika, by Toshi Maruki; When the Wind Blows, by Raymond Briggs. Bulletin of the Atomic Scientists, 39(2):39–40, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Messer:1983:BRB

[Mes83b] Robert L. Messer. Book review: Russian Roulette: The Superpower Game, by Arthur Macy Cox. Bulletin of the Atomic Scientists, 39(5):41–42, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Messer:1984:BRBb

[Mes84a] Robert L. Messer. Book review: Foreign Policy Choices for Americans: A Nonpartisan Guide for Voters. Bulletin of the Atomic Scientists, 40(9):52, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Messer:1984:BRBa

[Mes84b] Robert L. Messer. Book review: The Faces of Power: Constancy and Change in United States Foreign Policy from Truman to Reagan, by Seyon Brown. Bulletin of the Atomic Scientists, 40(6):38– 39, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Messer:1985:BRB

[Mes85a] Robert L. Messer. Book review: The Day the Sun Rose Twice, by Ferenc Morton Szasz. Bulletin of the Atomic Scientists, 41 (6):46–48, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Messer:1985:NET

[Mes85b] Robert L. Messer. New evidence on Truman’s decision. Bulletin of the Atomic Scientists, 41(7):50–56, August 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bak86].

Messer:1987:BRB

[Mes87] Robert L. Messer. Book review: The Making of the Atomic Bomb, by Richard Rhodes. Bulletin of the Atomic Scientists, 43(9):46– 47, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meselson:1991:MCS

[Mes91] Matthew Meselson. The myth of chemical superweapons. Bulletin of the Atomic Scientists, 47(3):12–15, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mesler:1997:LDU

[Mes97] Bill Mesler. Letter: Depleted uranium a long-term risk. Bulletin of the Atomic Scientists, 53(6):63–64, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark97e].

Metlay:1983:BRB

[Met83] Daniel Metlay. Book reviews: Radwaste: A Reporter’s Investigation of a Growing Nuclear Menace, by Fred C. Shapiro; The Politics of Nuclear Waste, edited by E. William Colglazier, Jr.; Radioactive Waste from Nuclear Power Plants, by Thomas Johanson and Peter Steen. Bulletin of the Atomic Scientists, 39(4): 34–35, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Metlay:2024:BRA

[Met24] Daniel Metlay. Becoming a responsible ancestor. Bulletin of the Atomic Scientists, 81(1):29–35, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2024.
2438494.

Meyer:1949:PRW

[Mey49a] Cord Meyer, Jr. A progress report on world federation. Bulletin of the Atomic Scientists, 5(10):281–282, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyerhoff:1949:NSF

[Mey49b] Howard A. Meyerhoff. National Science Foundation, 1949. Bulletin of the Atomic Scientists, 5(1):2, 32, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyerhoff:1954:AM

[Mey54] Howard A. Meyerhoff. Arms and manpower. Bulletin of the Atomic Scientists, 10(4):119–120, 122, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyerhoff:1956:PVS

[Mey56] Howard A. Meyerhoff. Part VI: Science, education, and communication: The plight of science education. Bulletin of the Atomic Scientists, 12(9):333–337, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyerhoff:1961:UP

[Mey61] Howard A. Meyerhoff. Useless publication. Bulletin of the Atomic Scientists, 17(3):92–94, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyer:1966:BC

[Mey66] Herbert M. Meyer. The beginning of the commonsense. Bulletin of the Atomic Scientists, 22(2):23–25, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyer:1969:CO

[Mey69] J. Meyer. Communications: The oppressions. Bulletin of the Atomic Scientists, 25(8):46, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyer:1979:MMM

[Mey79] Stephen M. Meyer. MAPS for the MX missile. Bulletin of the Atomic Scientists, 35(6):26–29, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyrowitz:1983:NWL

[Mey83] Elliott L. Meyrowitz. Are nuclear weapons legal? Bulletin of the Atomic Scientists, 39(8):49–52, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyrowitz:1985:NWI

[Mey85] Elliott L. Meyrowitz. Nuclear weapons are illegal threats. Bulletin of the Atomic Scientists, 41(5):35–37, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meyers:1986:LNT

[Mey86] Aaron L. Meyers. Letter: Nuclear testing. Bulletin of the Atomic Scientists, 42(7):66–67, August/September 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DM86].

Meyer:1994:GOR

[Mey94] Eugene L. Meyer. Guest opinion: Revisionism, revised. Bulletin of the Atomic Scientists, 50(6):4, November/December 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Loe95].

Meyers:2005:O

[Mey05] Deborah W. Meyers. The outsiders. Bulletin of the Atomic Scientists, 61(6):66–68, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/6/66.

Meyer:2008:RPS

[Mey08] John M. Meyer. Rethinking personal sacrifice. Bulletin of the Atomic Scientists, 64(5):??, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/5/12.

Meyer:2018:OWT

[Mey18] Paul Meyer. Our WMD treaties are working. Bulletin of the Atomic Scientists, 74(1):27–31, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2017.
1413058.

Mez:2012:GME

[Mez12] Lutz Mez. Germany’s merger of energy and climate change policy. Bulletin of the Atomic Scientists, 68(6):22–29, November/ December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/68/
6/22.

McNamara:1963:NRM

[MF63] Robert S. McNamara and Bernard T. Feld. News & report: Mc-Namara on strategy: A change in policy. Bulletin of the Atomic Scientists, 19(4):36–39, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mishra:1968:NEI

[MF68] Indu Shekhar Mishra and Bernard T. Feld. The NPT is not enough for India. Bulletin of the Atomic Scientists, 24(6):4–5, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maurer:2010:HCD

[MF10] Stephen M. Maurer and Markus Fischer. How to control dualuse technologies in the age of global commerce. Bulletin of the Atomic Scientists, 66(1):41–47, January 2010. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/66/1/41.

Marseille:1962:LEC

[MFC62] Walter W. Marseille, C. A. Favill, and George Cunningham. Letter to the Editor: Civil defense. Bulletin of the Atomic Scientists, 18(5):40–41, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Malygina:2017:AHT

[MFN17] Anastasia Malygina, Sven-Eric Fikenscher, and Jenny Nielsen. Amid high tensions, an urgent need for nuclear restraint. Bulletin of the Atomic Scientists, 73(4):279–283, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1338049.

Mian:2008:FNL

[MG08] Zia Mian and Alexander Glaser. A frightening nuclear legacy. Bulletin of the Atomic Scientists, 64(4):??, September 2008. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/64/4/42.

Miasnikov:2007:GS

[MGZH07] Eugene Miasnikov, Mark Gubrud, Robert Zubrin, and Regina Hagen. Governing space. Bulletin of the Atomic Scientists, 63 (5):18–19, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/5/18.

McDonnell:1967:CS

[MH67] R. A. McDonnell and Kurt H. Hohenmeier. Comments: The SST. Bulletin of the Atomic Scientists, 23(5):64, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Markusen:1984:SS

[MH84] Eric Markusen and John B. Harris. Social sciences. Bulletin of the Atomic Scientists, 40(10):10s–12s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendelsohn:1989:CBT

[MH89a] Jack Mendelsohn and Thomas Halverson. The conventional balance: a TKO for NATO? Bulletin of the Atomic Scientists, 45(2): 30–34, 36–40, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HFMH89].

Mendelsohn:1989:GHM

[MH89b] Jack Mendelsohn and Thomas Halverson. In the gift horse’s mouth. Bulletin of the Atomic Scientists, 45(2):33, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marty:2005:WYD

[MHD^{+}05] Martin E. Marty, Pervez Hoodbhoy, Thomas Donnelly, Robert L. Gallucci, Gar Alperovitz, Richard B. Frank, Mary Palevsky, and Tadatoshi Akiba. Would you have dropped the bomb? Bulletin of the Atomic Scientists, 61(4):50–63, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/4/50.

Mitrovica:2018:ASL

[MHKL18] Jerry X. Mitrovica, Carling Hay, Robert Kopp, and Megan Lickley. All sea level is local. Bulletin of the Atomic Scientists, 74 (3):142–147, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1461935.

Moss:1990:LWF

[MHM90] Norman Moss, Daniel Hirsch, and William G. Mathews. Letters: What Fuchs didn’t know, and when he didn’t know it. Bulletin of the Atomic Scientists, 46(4):51–52, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90a].

Mian:1998:LBN

[Mia98] Zia Mian. Letter: Bringing the NPT back into play. Bulletin of the Atomic Scientists, 54(3):67–68, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bla97].

Mian:2015:NSS

[Mia15a] Zia Mian. Out of the nuclear shadow: Scientists and the struggle against the Bomb. Bulletin of the Atomic Scientists, 71(1):59– 69, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miasnikov:2015:HAN

[Mia15b] Eugene Miasnikov. How to approach nuclear modernization? : a Russian response. Bulletin of the Atomic Scientists, 71(3):12–15, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mian:2016:SUS

[Mia16] Zia Mian. Should the United States begin talks to ban nuclear weapons? Bulletin of the Atomic Scientists, 72(5):280–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michiels:1951:BAS

[Mic51] J. H. Michiels. The British Atomic Scientists’ Association. Bulletin of the Atomic Scientists, 7(1):25–26, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michael:1955:CBU

[Mic55] Donald N. Michael. Civilian behavior under atomic bombardment. Bulletin of the Atomic Scientists, 11(5):173–177, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michael:1961:RRA

[Mic61] Donald N. Michael. Ritualized rationality and arms control. Bulletin of the Atomic Scientists, 17(2):71–73, February 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michael:1962:PNW

[Mic62] Donald N. Michael. Psychopathology of nuclear war. Bulletin of the Atomic Scientists, 18(5):28–29, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michaelis:1964:SI

[Mic64] Michael Michaelis. A strategy for innovation. Bulletin of the Atomic Scientists, 20(4):19–23, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See errata [Ano64k].

Michaelis:1968:CWB

[Mic68] Michael Michaelis. Can we build the world we want? Bulletin of the Atomic Scientists, 24(1):43–49, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michelson:1969:SAP

[Mic69] Irving Michelson. A Space Age phenomenon: The evolution of lunar studies. Bulletin of the Atomic Scientists, 25(7):52–55, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michaelis:1973:CD

[Mic73] Anthony R. Michaelis. Coping with disaster. Bulletin of the Atomic Scientists, 29(4):24–29, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Michelson:1974:TTT

[Mic74a] Irving Michelson. Tides’ tortured theory. Bulletin of the Atomic Scientists, 30(3):31–34, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Micklin:1974:EHN

[Mic74b] Philip P. Micklin. Environmental hazards of nuclear waste. Bulletin of the Atomic Scientists, 30(4):36–42, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hur74].

Micallef:1981:CNB

[Mic81] Joseph V. R. Micallef. Commentary: A nuclear bomb for Libya? Bulletin of the Atomic Scientists, 37(7):14–15, August/ September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Middleton:1965:MHH

[Mid65] John T. Middleton. Man and his habitat: Problems of pollution. Bulletin of the Atomic Scientists, 21(3):18–22, March 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Middlehurst:1973:ALP

[Mid73] Barbara M. Middlehurst. Active lunar processes: Moonquakes and transient events. Bulletin of the Atomic Scientists, 29(10):35– 41, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miettinen:1974:CAT

[Mie74] J. K. Miettinen. The chemical arsenal: The time to defuse is now. Bulletin of the Atomic Scientists, 30(7):37–43, September 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miefert:1976:LBP

[Mie76] John F. Miefert. Letter: ‘blatant propaganda’. Bulletin of the Atomic Scientists, 32(1):3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch75e].

Miettinen:1977:ERW

[Mie77] Jorma K. Miettinen. Enhanced radiation warfare. Bulletin of the Atomic Scientists, 33(7):32–37, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miettinen:1979:CFD

[Mie79] Jorma K. Miettinen. Commentary from Finland: Détente in 1979. Bulletin of the Atomic Scientists, 35(1):8–9, January 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1947:ASP

[Mil47] Merle Miller. The atomic scientists in politics. Bulletin of the Atomic Scientists, 3(9):242, 252, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1948:FOR

[Mil48a] Byron S. Miller. The first official report on AEC patent problems. Bulletin of the Atomic Scientists, 4(3):77–79, March 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1948:LPA

[Mil48b] Byron S. Miller. A law is passed — the Atomic Energy Act of 1946. The University of Chicago Law Review, 15(4):799–821, Summer 1948. ISSN 0041-9494 (print), 1939-859X (electronic).

URL http://www.jstor.org/stable/10.2307/1597968.

Miller:1950:ECI

[Mil50] Byron S. Miller. Easing controls over industrial development — a comment on the Lilienthal proposals. Bulletin of the Atomic Scientists, 6(8–9):279–282, 286, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Miller:1952:AEA

[Mil52] Byron S. Miller. Atomic Energy Act: Second stage. Bulletin of the Atomic Scientists, 8(1):14–16, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miles:1962:LEA

[Mil62] Charles P. Miles. Letter to the Editor: Asks fair shake for CD. Bulletin of the Atomic Scientists, 18(1):37–38, January 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1963:LES

[Mil63a] Arthur S. Miller. Letter to the Editor: Space research. Bulletin of the Atomic Scientists, 19(1):32, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Millis:1963:LES

[Mil63b] Walter Millis. Letter to the Editor: Strategy. Bulletin of the Atomic Scientists, 19(1):32–33, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1968:MWS

[Mil68] Raphael Miller. The metaphysical world of strategic systems. Bulletin of the Atomic Scientists, 24(10):18–22, December 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1969:BBC

[Mil69a] Arthur Selwyn Miller. Books: The Closed Corporation: American Universities in Crisis, by James Ridgeway. Bulletin of the Atomic Scientists, 25(4):45–46, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1969:RTC

[Mil69b] Arthur Selwyn Miller. The rise of the techno-corporate state in America. Bulletin of the Atomic Scientists, 25(1):14–19, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1969:CDS

[Mil69c] Paul R. Miller. The Chicago demonstrations: A study in identity. Bulletin of the Atomic Scientists, 25(4):3–6, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1970:RAC

[Mil70] Paul R. Miller. Revolutionists among the Chicago demonstrators. Bulletin of the Atomic Scientists, 26(2):16–21, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1971:GEC

[Mil71] Edward F. Miller, Jr. Gas–electric car. Bulletin of the Atomic Scientists, 27(2):2, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SEM70].

Mills:1978:CLT

[Mil78] Craig M. Mills. Commentary: Later than you think. Bulletin of the Atomic Scientists, 34(10):57, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1979:HNW

[Mil79] George H. Miller. How a nuclear war would be fought … at sea. Bulletin of the Atomic Scientists, 35(5):25–26, May 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milstein:1981:SRP

[Mil81] Mikhail A. Milstein. Special report from Paris: Science, technology and the arms buildup — II. Bulletin of the Atomic Scientists, 37(6):55–59, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1983:CBW

[Mil83a] Richard B. Miller. Catholic bishops on war. Bulletin of the Atomic Scientists, 39(5):9–13, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1983:CFD

[Mil83b] Richard B. Miller. Commentary: First draft of the pastoral letter. Bulletin of the Atomic Scientists, 39(7):55–56, August/ September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1984:CFB

[Mil84] Richard B. Miller. Commentary: The French bishops on deterrence. Bulletin of the Atomic Scientists, 40(5):58–59, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milhollin:1985:BRB

[Mil85a] Gary Milhollin. Book review: The Plutonium Business and the Spread of the Bomb, by Walter C. Patterson. Bulletin of the Atomic Scientists, 41(11):61–62, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1985:BRBb

[Mil85b] Richard Miller. Book review: Can Modern War Be Just?, by James Turner Johnson; Moral Principles and Nuclear Weapons, by Douglas P. Lackey. Bulletin of the Atomic Scientists, 41(5): 45–47, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [JM85].

Miller:1985:BRBa

[Mil85c] Steven E. Miller. Book review: The Defense Reform Debate, by Asa A. Clark IV, Peter W. Chiarelli, Jeffrey S. McKitrick, and James W. Reed. Bulletin of the Atomic Scientists, 41(3):54, 56, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1985:TW

[Mil85d] Steven E. Miller. Technology and war. Bulletin of the Atomic Scientists, 41(11):46–48, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1986:UCD

[Mil86a] Richard L. (Richard Lee) Miller. Under the cloud: the decades of nuclear testing. Free Press, New York, NY, USA, 1986. ISBN 0-02-921620-6. xii + 547 + 8 pp. LCCN U264 .M55 1986.

Miller:1986:BRB

[Mil86b] Steven E. Miller. Book reviews: The Pentagon and the Art of War, by Edward N. Luttwak; Toward a More Effective Defense, edited by Barry M. Blechman and William J. Lynn. Bulletin of the Atomic Scientists, 42(5):41–43, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milliken:1987:ANG

[Mil87] Robert Milliken. Australia’s nuclear graveyard. Bulletin of the Atomic Scientists, 43(3):38–44, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Millar:1988:BRB

[Mil88] James R. Millar. Book review: Gorbachev’s Challenge: Economics Reform in the Age of High Technology, by Marshall Goldman. Bulletin of the Atomic Scientists, 44(3):52, 54, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milhollin:1989:IML

[Mil89a] Gary Milhollin. India’s missiles — with a little help from our friends. Bulletin of the Atomic Scientists, 45(9):31–35, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1989:LBT

[Mil89b] Richard Miller. Letter: Buy tritium from Moscow. Bulletin of the Atomic Scientists, 45(3):44, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1991:LRM

[Mil91] Clark A. Miller. Letter: Review missed target. Bulletin of the Atomic Scientists, 47(10):43–44, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pit91].

Milner:1995:LOS

[Mil95] Paul F. Milner. Letter: The other side. Bulletin of the Atomic Scientists, 51(2):70, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milner:2002:TSV

[Mil02] Glen Milner. Trident: Subs v. the environment. Bulletin of the Atomic Scientists, 58(4):11–13, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/4/11.

Miller:2003:RDT

[Mil03] Eric Miller. The real “Don’t Tell”. Bulletin of the Atomic Scientists, 59(4):73, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/4/73.

Miller:2010:SSB

[Mil10] Jerry Miller. Stockpile: the story behind 10,000 strategic nuclear weapons. Naval Institute Press, Annapolis, MD, USA, 2010. ISBN 1-59114-531-7. xix + 273 pp. LCCN UA23 .M5639 2010.

Miller:2012:NWM

[Mil12a] Steven E. Miller. Nuclear weapons 2011: Momentum slows, reality returns. Bulletin of the Atomic Scientists, 68(1):20–28, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/1/20.

Mills:2012:WRC

[Mil12b] Evan Mills. Weighing the risks of climate change mitigation strategies. Bulletin of the Atomic Scientists, 68(6):67–78, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/6/67.

Miller:2020:NNF

[Mil20] James N. Miller. No to no first use — for now. Bulletin of the Atomic Scientists, 76(1):8–13, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2019.
1701278.

Minnaert:1948:OAE

[Min48] M. Minnaert. Opinions on atomic energy from The Netherlands. Bulletin of the Atomic Scientists, 4(3):87, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Minnaert:1952:SOE

[Min52] M. Minnaert. Some other European experiences. Bulletin of the Atomic Scientists, 8(7):249, 261, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Minor:1955:PMP

[Min55] John R. Minor. Political machinery for peace. Bulletin of the Atomic Scientists, 11(3):85–86, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Minton:1986:LNF

[Min86] Charles Minton. Letter: Nuclear fear. Bulletin of the Atomic Scientists, 42(5):46, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cal86a].

Minton:1988:LMF

[Min88] Charles Minton. Letter: Misunderstood freeze. Bulletin of the Atomic Scientists, 44(8):54, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LC88].

Minutaglio:1994:BCW

[Min94] Bill Minutaglio. Boley Caldwell wants an apology. Bulletin of the Atomic Scientists, 50(3):35–38, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miramontes:1989:WMN

[Mir89] Octavio Miramontes. Wooing Mexico to nuclear power. Bulletin of the Atomic Scientists, 45(6):36–38, July/August 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mitchell:1950:PFK

[Mit50] Clyde Mitchell. Promises and facts in Korea and China. Bulletin of the Atomic Scientists, 6(12):357–385, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mitchell:1954:BBC

[Mit54] C. Clyde Mitchell. Books: The Challenge of Man’s Future, by Harrison Brown. Bulletin of the Atomic Scientists, 10(8):329– 330, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mitra:1958:LEI

[Mit58] A. K. Mitra. Letter to the Editor: Indian plan for banning of nuclear tests. Bulletin of the Atomic Scientists, 14(2):96, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Fry58].

Mitchell:1962:SAE

[Mit62] Lawrence C. Mitchell. Soviet–American exchange of scientists. Bulletin of the Atomic Scientists, 18(2):15–17, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mitchell:1970:AWS

[Mit70] Charles I. Mitchell. Los Alamos: From weapon shop to scientific laboratory. Bulletin of the Atomic Scientists, 26(9):24–27, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mittelman:1983:COW

[Mit83] James H. Mittelman. Commentary: The other war in the South Atlantic. Bulletin of the Atomic Scientists, 39(4):38–39, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mitchell:1997:ASD

[Mit97] Gordon Mitchell. Another strategic deception initiative. Bulletin of the Atomic Scientists, 53(2):22–23, March/April 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://www.highbeam.com/doc/1G1-19201489.html.

Mansson:1968:TPU

[MJ68] Helge Hilding Mansson and Walter Johnson. Trouble in paradise: The University of Hawaii and Dr. Oliver Lee. Bulletin of the Atomic Scientists, 24(7):10–16, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moskos:1980:CMA

[MJ80] Charles C. Moskos and Morris Janowitz. Commentary: Making the all-volunteer army work. Bulletin of the Atomic Scientists, 36(6):6, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mitchell:1981:CPP

[MJ81] Theodore R. Mitchell and Giff Johnson. Commentary: In protest of a ‘Paradise Lost’. Bulletin of the Atomic Scientists, 37(4):

55–57, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Joh80b].

Myers:1992:LOH

[MJD92] R. Thomas Myers, John Jirikowic, and Paul E. Damon. Letter: On the other hand …. Bulletin of the Atomic Scientists, 48 (7):45–46, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sin92].

Metzner:1951:WPT

[MK51] Charles A. Metzner and Julia B. Kessler. What are the people thinking? Bulletin of the Atomic Scientists, 7(11):341, 352, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mendlovitz:1979:CIP

[MK79] Saul Mendlovitz and Rajni Kothari. Commentary: An indictment: The perversion of science and technology. Bulletin of the Atomic Scientists, 35(1):57–59, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1980:CCR

[MK80] John Martin and George W. Kneale. Commentary: On cancer and radiation. Bulletin of the Atomic Scientists, 36(9):59–60, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maroncelli:2002:TGN

[MK02a] James M. Maroncelli and Timothy L. Karpin. The traveler’s guide to nuclear weapons: a journey through America’s cold war battlefields. Historical Odysseys Pub., Silverdale, WA, USA, 2002. ISBN 0-9725051-0-5. ???? pp. LCCN U264.3. One CD-ROM.

Musalo:2002:UP

[MK02b] Karen Musalo and Stephen Knight. Unequal protection. Bulletin of the Atomic Scientists, 58(6):56–61, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/6/56.

Maroncelli:2003:YTS

[MK03] James M. Maroncelli and Timothy L. Karpin. Not your typical snapshots. Bulletin of the Atomic Scientists, 59(6):72–73, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/6/72.

Margen:1975:CNP

[ML75] Peter Margen and Sören Lindhe. The capacity of nuclear power plants. Bulletin of the Atomic Scientists, 31(8):38–40, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com74, Com75h] and reply [Com75f].

Myrdal:1982:NNW

[ML82] Alva Myrdal and Milton Leitenberg. On a Nordic nuclearweapon-free zone. Bulletin of the Atomic Scientists, 38(7):61– 63, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei82a].

Morton:1988:EEA

[ML88] Brian Morton and Joanne Landy. East European activists test glasnost. Bulletin of the Atomic Scientists, 44(4):18–24, 26, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wol88].

Mowatt-Larssen:2009:ATP

[ML09] Rolf Mowatt-Larssen. The Armageddon Test: Preventing nuclear terrorism. Bulletin of the Atomic Scientists, 65(5):60–70, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/5/60.

Mowatt-Larssen:2010:NNT

[ML10] Rolf Mowatt-Larssen. Nightmares of nuclear terrorism. Bulletin of the Atomic Scientists, 66(2):37–45, March 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/66/2/37.

Monson:1951:PUD

[MM51] Donald Monson and Astrid Monson. A program for urban dispersal. Bulletin of the Atomic Scientists, 7(9):244–250, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Meinel:1971:ITN

[MM71] Aden Baker Meinel and Marjorie Pettit Meinel. Is it time for a new look at solar energy? Bulletin of the Atomic Scientists, 27(8):32–37, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marcy:1977:LPR

[MM77] Carl Marcy and Jeanne Mattison. Letter: Pressure from the right. Bulletin of the Atomic Scientists, 33(6):2, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel77e].

Mann:1982:CAP

[MM82] Robert Mann and Gerald E. Marsh. Commentary: On ‘Atoms for Peace’. Bulletin of the Atomic Scientists, 38(8):58, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar82c].

Mott:1986:PAP

[MM86] Frank L. Mott and Susan H. Mott. Population: The African paradox of growth and development. Bulletin of the Atomic Scientists, 42(4):26–29, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Miller:1990:GWN

[MM90] Alan Miller and Irving Mintzer. Global warming: No nuclear quick fix. Bulletin of the Atomic Scientists, 46(5):30–32, 34, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Massie:1998:NTB

[MM98] Suzanne Massie and Priscilla McMillan. NATO: Turn back now. Bulletin of the Atomic Scientists, 54(1):38–39, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mazurana:2001:CSW

[MM01] Dyan Mazurana and Susan McKay. Child soldiers; what about the girls? Bulletin of the Atomic Scientists, 57(5):30–35, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/5/30.

Mousseau:2011:LPL

[MM11] Timothy A. Mousseau and Anders P. Møller. Landscape portrait: a look at the impacts of radioactive contaminants on Chernobyl’s wildlife. Bulletin of the Atomic Scientists, 67(2):38–46, March/ April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/67/2/38.

Mian:1996:NST

[MN96] Zia Mian and A. H. Nayyar. The nuclear subcontinent: A time of testing? Bulletin of the Atomic Scientists, 52(4):35–40, July/ August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshak:1946:ABD

[MNS46] R. E. Marshak, E. C. Nelson, and L. I. Schiff. Atomic bomb damage — Japan and USA. Bulletin of the Atomic Scientists, 1 (10):6–7, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moch:1956:TF

[Moc56] Jules Moch. Technology and the future. Bulletin of the Atomic Scientists, 12(4):112–118, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moglewer:1981:ISN

[Mog81] Sidney Moglewer. IAEA safeguards and non-proliferation. Bulletin of the Atomic Scientists, 37(8):24–29, October 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [FM82, Fis82] and addendum [Ano82a].

Moldauer:1962:BBE

[Mol62a] Peter A. Moldauer. Books: The Effects of Nuclear Weapons, revised edition, edited by Samuel Glasstone. Bulletin of the Atomic Scientists, 18(9):26–27, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moldauer:1962:PS

[Mol62b] Peter A. Moldauer. Population survival. Bulletin of the Atomic Scientists, 18(5):19–24, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moldauer:1969:ACT

[Mol69] Peter Moldauer. The ABM comes to town. Bulletin of the Atomic Scientists, 25(1):4–6, 20, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moler:1989:LCR

[Mol89] Robert B. Moler. Letter: Circular reasoning? Bulletin of the Atomic Scientists, 45(5):47, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet89a].

Moller:1990:LCN

[Møl90] Bjørn Møller. Letter: Conventional, nuclear wars different disasters. Bulletin of the Atomic Scientists, 46(2):45–46, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SS89].

Molinelli:1996:LCL

[Mol96] Sonny Molinelli. Letter: The Chicken Little syndrome. Bulletin of the Atomic Scientists, 52(2):3, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo96g].

Moltz:2003:RSC

[Mol03] James Clay Moltz. Reining in the space cowboys. Bulletin of the Atomic Scientists, 59(1):61–66, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/1/61.

Monod:1952:SFE

[Mon52] Jacques Monod. Some French experiences. Bulletin of the Atomic Scientists, 8(7):236, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Monod:1953:LEE

[Mon53a] Jacques Monod. Letter to the Editor: the execution of Ethel and Julius Rosenberg. Bulletin of the Atomic Scientists, 9(8):319– 320, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See letter [Mar53c].

Monson:1953:CPP

[Mon53b] Donald Monson. City planning in Project East River. Bulletin of the Atomic Scientists, 9(7):265–267, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Monson:1954:DO

[Mon54] Donald Monson. Is dispersal obsolete? Bulletin of the Atomic Scientists, 10(10):378–383, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Monroe:1984:CWA

[Mon84] Manus Monroe. Commentary: Wisdom of the ages. Bulletin of the Atomic Scientists, 40(5):60, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Monfort:1989:ASW

[Mon89] Charles A. Monfort. ASATs: Star Wars on the cheap. Bulletin of the Atomic Scientists, 45(3):10–13, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1951:DLG

[Moo51] Lyman S. Moore. Democratic local government in the AEC communities. Bulletin of the Atomic Scientists, 7(1):19–22, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:CP

[Moo91a] Mike Moore. Chasing pirates. Bulletin of the Atomic Scientists, 47(8):28, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:EDD

[Moo91b] Mike Moore. From the Editors: A distant drum. Bulletin of the Atomic Scientists, 47(10):2, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HJS^{+}92, HD92, Pet92].

Moore:1991:EJH

[Moo91c] Mike Moore. From the Editors: A journal of hope. Bulletin of the Atomic Scientists, 47(6):2, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:ETP

[Moo91d] Mike Moore. From the Editors: A time for prodding. Bulletin of the Atomic Scientists, 47(9):2, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:EIC

[Moo91e] Mike Moore. From the Editors: Iraq, continued. Bulletin of the Atomic Scientists, 47(7):2, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:ETM

[Moo91f] Mike Moore. From the Editors: Timely matters. Bulletin of the Atomic Scientists, 47(8):2, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:HGB

[Moo91g] Mike Moore. How George Bush won his spurs. Bulletin of the Atomic Scientists, 47(8):26–33, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1991:TH

[Moo91h] Mike Moore. A “toy handcuff’. Bulletin of the Atomic Scientists, 47(8):32, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:SSW

[Moo92a] Mike Moore. As surely as the sun will rise. Bulletin of the Atomic Scientists, 48(6):25, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENB

[Moo92b] Mike Moore. Editor’s note: Brilliant but dumb. Bulletin of the Atomic Scientists, 48(3):2, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENCc

[Moo92c] Mike Moore. Editor’s note: Casting stones. Bulletin of the Atomic Scientists, 48(6):2, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENCb

[Moo92d] Mike Moore. Editor’s note: Childhood’s end. Bulletin of the Atomic Scientists, 48(4):2, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENCa

[Moo92e] Mike Moore. Editor’s note: Cold War victory. Bulletin of the Atomic Scientists, 48(2):2, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENCd

[Moo92f] Mike Moore. Editor’s note: Consequences. Bulletin of the Atomic Scientists, 48(10):2, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:END

[Moo92g] Mike Moore. Editor’s note: Do the right thing. Bulletin of the Atomic Scientists, 48(5):2, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENNb

[Moo92h] Mike Moore. Editor’s note: National insecurity. Bulletin of the Atomic Scientists, 48(9):2, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENNa

[Moo92i] Mike Moore. Editor’s note: Neighbors. Bulletin of the Atomic Scientists, 48(8):2, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ENU

[Moo92j] Mike Moore. Editor’s note: The untried path. Bulletin of the Atomic Scientists, 48(7):2, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:EPC

[Moo92k] Mike Moore. From the Editors: A precious chance. Bulletin of the Atomic Scientists, 48(1):2, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1992:ISF

[Moo92l] Mike Moore. The incident at Stagg Field. Bulletin of the Atomic Scientists, 48(10):11–15, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). 50th anniversary of the first controlled nuclear chain reaction at the University of Chicago, 2 December 1942. See comment [Phi93].

Moore:1993:BRN

[Moo93a] Mike Moore. Book review: The new war: The Once and Future Superpower, by Joseph J. Romm. Bulletin of the Atomic Scientists, 49(3):44–45, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:BRZ

[Moo93b] Mike Moore. Book review: The zero option: A Nuclear-Weapon-Free World, edited by Joseph Rotblat, Jack Steinberger, and Bhalchandra Udgaonkar. Bulletin of the Atomic Scientists, 49 (10):48–49, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [KS94, Phi94].

Moore:1993:ENA

[Moo93c] Mike Moore. Editor’s note: Apocalypse then. Bulletin of the Atomic Scientists, 49(3):2, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENB

[Moo93d] Mike Moore. Editor’s note: Beyond the yurt. Bulletin of the Atomic Scientists, 49(8):2, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENE

[Moo93e] Mike Moore. Editor’s note: Eating grass. Bulletin of the Atomic Scientists, 49(5):2, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENF

[Moo93f] Mike Moore. Editor’s note: Felony murder. Bulletin of the Atomic Scientists, 49(6):2, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENM

[Moo93g] Mike Moore. Editor’s note: Messy, but useful. Bulletin of the Atomic Scientists, 49(9):2, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENN

[Moo93h] Mike Moore. Editor’s note: New imperatives. Bulletin of the Atomic Scientists, 49(2):2, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENR

[Moo93i] Mike Moore. Editor’s note: Reasons of state. Bulletin of the Atomic Scientists, 49(4):2, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENTa

[Moo93j] Mike Moore. Editor’s note: Too much SYFS. Bulletin of the Atomic Scientists, 49(7):2, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1993:ENTb

[Moo93k] Mike Moore. Editor’s note: Trust me. Bulletin of the Atomic Scientists, 49(10):2, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:BRW

[Moo94a] Mike Moore. Book review: White crows: The Making of a Soviet Scientist, by Roald Z. Sagdeev. Bulletin of the Atomic Scientists, 50(5):56–57, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:BAC

[Moo94b] Mike Moore. Bulletin: Avner Cohen, meet Franz Kafka. Bulletin of the Atomic Scientists, 50(5):5–6, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:ENCa

[Moo94c] Mike Moore. Editor’s note: Change and rebirth. Bulletin of the Atomic Scientists, 50(1):2, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:ENCb

[Moo94d] Mike Moore. Editor’s note: Crossroads. Bulletin of the Atomic Scientists, 50(3):2, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:ENG

[Moo94e] Mike Moore. Editor’s note: God’s gift. Bulletin of the Atomic Scientists, 50(6):2, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bra95].

Moore:1994:ENH

[Moo94f] Mike Moore. Editor’s note: Hard questions. Bulletin of the Atomic Scientists, 50(2):2, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:ENL

[Moo94g] Mike Moore. Editor’s note: Lying well. Bulletin of the Atomic Scientists, 50(4):2, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:ENN

[Moo94h] Mike Moore. Editor’s note: No-brainer. Bulletin of the Atomic Scientists, 50(5):2, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:IAB

[Moo94i] Mike Moore. Introduction: The able-baker-where’s-Charlie follies. Bulletin of the Atomic Scientists, 50(3):24–26, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1994:ISS

[Moo94j] Mike Moore. Introduction to Sarajevo stories. Bulletin of the Atomic Scientists, 50(2):30–32, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:ABG

[Moo95a] Mike Moore. Alpha, beta, gamma. Bulletin of the Atomic Scientists, 51(1):58, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:BRI

[Moo95b] Mike Moore. Book review: Imaginary enemies: Rogue States and Nuclear Outlaws: America’s Search for a New Foreign Policy, by Michael Klare. Bulletin of the Atomic Scientists, 51(5):56–57, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:BRJ

[Moo95c] Mike Moore. Book review: Just business: And Weapons for All, by William D. Hartung. Bulletin of the Atomic Scientists, 51(2):

65–67, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:BRG

[Moo95d] Mike Moore. Book review: The gadgets made real: Picturing the Bomb: Photographs from the Secret World of the Manhattan Project, by Rachel Fermi and Esther Samra, introduction by Richard Rhodes. Bulletin of the Atomic Scientists, 51(6):65–66, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-17459604.html. See [FS95].

Moore:1995:BRT

[Moo95e] Mike Moore. Book review: Troublesome imagery: Hiroshima in America: Fifty Years of Denial, by Robert Jay Lifton and Greg Mitchell. Bulletin of the Atomic Scientists, 51(4):73–74, July/ August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WB95].

Moore:1995:ENC

[Moo95f] Mike Moore. Editor’s note: A counterweight. Bulletin of the Atomic Scientists, 51(5):2, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:ENF

[Moo95g] Mike Moore. Editor’s note: A fine madness. Bulletin of the Atomic Scientists, 51(6):2, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:END

[Moo95h] Mike Moore. Editor’s note: Dehousing. Bulletin of the Atomic Scientists, 51(3):2, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CMM95].

Moore:1995:ENM

[Moo95i] Mike Moore. Editor’s note: Make it so. Bulletin of the Atomic Scientists, 51(2):2, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:ENT

[Moo95j] Mike Moore. Editor’s note: Tight little island. Bulletin of the Atomic Scientists, 51(4):2, July/August 1995. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Her95b].

Moore:1995:ENV

[Moo95k] Mike Moore. Editor’s note: Visionaries. Bulletin of the Atomic Scientists, 51(1):2, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:MWJ

[Moo95l] Mike Moore. In memoriam: Walter J. Blum. Bulletin of the Atomic Scientists, 51(2):9, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:MNC

[Moo95m] Mike Moore. Midnight never came. Bulletin of the Atomic Scientists, 51(6):16–27, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:MSL

[Moo95n] Mike Moore. More security for less money. Bulletin of the Atomic Scientists, 51(5):34–37, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:WPD

[Moo95o] Mike Moore. So where’s the peace dividend? Bulletin of the Atomic Scientists, 51(5):30–31, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Toy96].

Moore:1995:TMI

[Moo95p] Mike Moore. Too much isn’t enough. Bulletin of the Atomic Scientists, 51(1):59, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:UPG

[Moo95q] Mike Moore. U.N. peacekeeping, a glass half empty, half full. Bulletin of the Atomic Scientists, 51(2):22–23, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1995:WS

[Moo95r] Mike Moore. A wonderful sight. Bulletin of the Atomic Scientists, 51(1):56, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:BCM

[Moo96a] Mike Moore. Behind the clock move. Bulletin of the Atomic Scientists, 52(2):17–23, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:BRD

[Moo96b] Mike Moore. Book review: Don’t panic: Revolution and War, by Stephen M. Walt. Bulletin of the Atomic Scientists, 52(5):60–61, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:BRP

[Moo96c] Mike Moore. Book review: Paradise delayed: The First Nuclear Era, by Alvin M. Weinberg. Bulletin of the Atomic Scientists, 52(1):55–56, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rau96].

Moore:1996:BA

[Moo96d] Mike Moore. A boost for abolition. Bulletin of the Atomic Scientists, 52(6):38–43, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:BWW

[Moo96e] Mike Moore. Bulletin: Wanted: wee weaponeers. Bulletin of the Atomic Scientists, 52(4):8–9, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:ENU

[Moo96f] Mike Moore. Editor’s note: An unassuming people. Bulletin of the Atomic Scientists, 52(4):2, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:ENSa

[Moo96g] Mike Moore. Editor’s note: On the scale. Bulletin of the Atomic Scientists, 52(1):2, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mol96].

Moore:1996:ENOa

[Moo96h] Mike Moore. Editor’s note: One sunny day. Bulletin of the Atomic Scientists, 52(5):2, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:ENOb

[Moo96i] Mike Moore. Editor’s note: Options. Bulletin of the Atomic Scientists, 52(6):2, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jon97].

Moore:1996:ENSb

[Moo96j] Mike Moore. Editor’s note: Shackling the genie. Bulletin of the Atomic Scientists, 52(2):2, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:ENM

[Moo96k] Mike Moore. Editor’s note: The moral principle. Bulletin of the Atomic Scientists, 52(3):2, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:KRP

[Moo96l] Mike Moore. Korean reactor pact on target. Bulletin of the Atomic Scientists, 52(2):8–9, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:RNS

[Moo96m] Mike Moore. Report: Nuclear summit: Cheery words, little action. Bulletin of the Atomic Scientists, 52(4):14–15, 60, July/ August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1996:WCS

[Moo96n] Mike Moore. World Court says mostly no to nuclear weapons. Bulletin of the Atomic Scientists, 52(5):39–42, September/ October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:BRR

[Moo97a] Mike Moore. Book review: Rationalizing the irrational: Nuclear Rites: A Weapons Laboratory at the End of the Cold War, by Hugh Gusterson. Bulletin of the Atomic Scientists, 53(4):56–57, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:BRS

[Moo97b] Mike Moore. Book review: Sink the SIOP: Caging the Nuclear Genie: An American Challenge for Global Security, by Stansfield Turner. Bulletin of the Atomic Scientists, 53(5):57–58, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:BDA

[Moo97c] Mike Moore. Bulletin: Dial up the abolition web site. Bulletin of the Atomic Scientists, 53(4):7, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:BTN

[Moo97d] Mike Moore. Bulletin: Teraflopping into the new millennium. Bulletin of the Atomic Scientists, 53(2):5–6, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:ENN

[Moo97e] Mike Moore. Editor’s note: A new direction. Bulletin of the Atomic Scientists, 53(3):2, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:ENA

[Moo97f] Mike Moore. Editor’s note: Abolition web. Bulletin of the Atomic Scientists, 53(2):2, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:ENC

[Moo97g] Mike Moore. Editor’s note: Commitments. Bulletin of the Atomic Scientists, 53(5):2, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:ENE

[Moo97h] Mike Moore. Editor’s note: Elasticity. Bulletin of the Atomic Scientists, 53(6):2, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:ENPb

[Moo97i] Mike Moore. Editor’s note: Pugwash at 40. Bulletin of the Atomic Scientists, 53(4):2, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:ENPa

[Moo97j] Mike Moore. Editor’s note: The Pacific century. Bulletin of the Atomic Scientists, 53(1):2, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:FYP

[Moo97k] Mike Moore. Forty years of Pugwash. Bulletin of the Atomic Scientists, 53(6):40–45, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DH98].

Moore:1997:LMB

[Moo97l] Mike Moore. Land mine ban close. Bulletin of the Atomic Scientists, 53(5):33, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:PDD

[Moo97m] Mike Moore. Plutonium: The disposal decision. Bulletin of the Atomic Scientists, 53(2):40–41, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:PT

[Moo97n] Mike Moore. Pu by the ton. Bulletin of the Atomic Scientists, 53(2):44, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:RND

[Moo97o] Mike Moore. Report: Nuclear disarmament: The NAS blueprint. Bulletin of the Atomic Scientists, 53(5):14–17, September/ October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:RTB

[Moo97p] Mike Moore. Report: Test ban treaty: Angles and loopholes. Bulletin of the Atomic Scientists, 53(4):11–13, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1997:TP

[Moo97q] Mike Moore. Taking the pledge. Bulletin of the Atomic Scientists, 53(6):43–45, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:ENH

[Moo98a] M. Moore. Editor’s note: Hopeful trends. Bulletin of the Atomic Scientists, 54(6):2, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:BRB

[Moo98b] Mike Moore. Book review: Nuclear Weapons in a Transformed World, edited by Michael J. Mazarr. Bulletin of the Atomic Scientists, 54(4):60–62, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:CP

[Moo98c] Mike Moore. Check, please! Bulletin of the Atomic Scientists, 54(5):34–35, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:ENC

[Moo98d] Mike Moore. Editor’s note: Chilling words. Bulletin of the Atomic Scientists, 54(4):2, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:END

[Moo98e] Mike Moore. Editor’s note: Dropping the ball. Bulletin of the Atomic Scientists, 54(2):2, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:ENM

[Moo98f] Mike Moore. Editor’s note: Moving on. Bulletin of the Atomic Scientists, 54(3):2, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:ENN

[Moo98g] Mike Moore. Editor’s note: Nine minutes. Bulletin of the Atomic Scientists, 54(5):2, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1998:ENW

[Moo98h] Mike Moore. Editor’s note: White House activist. Bulletin of the Atomic Scientists, 54(1):2–3, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1999:CCH

[Moo99a] Mike Moore. Call China’s hand. Bulletin of the Atomic Scientists, 55(4):21–23, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/4/21.

Moore:1999:CKC

[Moo99b] Mike Moore. China: The Kafkaesque case of Hua Di. Bulletin of the Atomic Scientists, 55(6):12–14, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/6/12.

Moore:1999:D

[Moo99c] Mike Moore. Demystification. Bulletin of the Atomic Scientists, 55(6):2, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/6/2.

Moore:1999:EK

[Moo99d] Mike Moore. Easy killing. Bulletin of the Atomic Scientists, 55(1):2, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/55/1/2.

Moore:1999:MHK

[Moo99e] Mike Moore. In memoriam: Henry Kendall, Glenn Seaborg, Gerhard Herzberg. Bulletin of the Atomic Scientists, 55(3):21–22, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
3/21.

Moore:1999:IW

[Moo99f] Mike Moore. Invisible war. Bulletin of the Atomic Scientists, 55(4):2, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/4/2.

Moore:1999:LMR

[Moo99g] Mike Moore. Leonard M. Rieser, 1922–1998. Bulletin of the Atomic Scientists, 55(2):4, March/April 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:1999:SP

[Moo99h] Mike Moore. A small project. Bulletin of the Atomic Scientists, 55(3):2, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/3/2.

Moore:1999:S

[Moo99i] Mike Moore. Splitism. Bulletin of the Atomic Scientists, 55(5): 2, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/5/2.

Moore:2000:BRB

[Moo00a] Mike Moore. Book review: Atomic Fragments: A Daughter’s Questions, by Mary Palevsky. Bulletin of the Atomic Scientists, 56(9):74–76, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

www.highbeam.com/doc/1G1-67151470.html.

Moore:2000:CJ

[Moo00b] Mike Moore. Con job? Bulletin of the Atomic Scientists, 56(2): 2, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
2/2.

Moore:2000:JA

[Moo00c] Mike Moore. Job application? Bulletin of the Atomic Scientists, 56(5):63–65, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/5/63.

Moore:2000:MIA

[Moo00d] Mike Moore. The meaning of it all. Bulletin of the Atomic Scientists, 56(6):75–76, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/6/75.

Moore:2000:NE

[Moo00e] Mike Moore. A new Editor. Bulletin of the Atomic Scientists, 56(3):2, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/3/2.

Moore:2000:Q

[Moo00f] Mike Moore. A question. Bulletin of the Atomic Scientists, 56(1): 2, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/1/2.

Moore:2000:RJS

[Moo00g] Mike Moore. Remembering John Simpson. Bulletin of the Atomic Scientists, 56(6):4–6, November/December 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/6/4.

Moore:2001:I

[Moo01a] Mike Moore. Interoperability. Bulletin of the Atomic Scientists, 57(5):68–70, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/5/68.

Moore:2001:SC

[Moo01b] Mike Moore. Space cop. Bulletin of the Atomic Scientists, 57 (3):68–70, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/3/68.

Moore:2001:SNA

[Moo01c] Mike Moore. Space: Non-aggressive weapons? Bulletin of the Atomic Scientists, 57(2):17–19, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/2/17.

Moore:2001:WSC

[Moo01d] Mike Moore. Watch out for space command. Bulletin of the Atomic Scientists, 57(1):24–25, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/24.

Moore:2001:BB

[Moo01e] Richard Moore. Bomb buddies. Bulletin of the Atomic Scientists, 57(3):67–68, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/3/67.

Moore:2001:WHM

[Moo01f] Richard Moore. Where her Majesty’s weapons were. Bulletin of the Atomic Scientists, 57(1):58–64, January/February 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/1/58.

Moore:2002:LB

[Moo02a] LeRoy Moore. Lowering the bar. Bulletin of the Atomic Scientists, 58(3):28–37, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/3/28.

Moore:2002:GC

[Moo02b] Mike Moore. A game of chance. Bulletin of the Atomic Scientists, 58(1):67–69, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/1/67.

Moore:2002:MDR

[Moo02c] Mike Moore. Missile defenses, relabeled. Bulletin of the Atomic Scientists, 58(4):22–23, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/22.

Moore:2002:TTB

[Moo02d] Mike Moore. Twinkle, twinkle: Book review: Sputnik: The Shock of the Century, by Paul Dickson Walker. Bulletin of the Atomic Scientists, 58(3):65–67, May/June 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/3/65;   http://
www.highbeam.com/doc/1G1-86128161.html.

Moore:2003:ACT

[Moo03a] Mike Moore. Article commentary: Truman got it right. Bulletin of the Atomic Scientists, 59(1):20–22, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/1/20.

Moore:2003:SCC

[Moo03b] Mike Moore. Space cops: Coming to a planet near you! Bulletin of the Atomic Scientists, 59(6):46–56, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/6/46.

Moore:2004:BHA

[Moo04] Mike Moore. Black hats all around. Bulletin of the Atomic Scientists, 60(3):62–65, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/3/62.

Mooney:2005:RO

[Moo05a] Chris Mooney. Requiem for an office. Bulletin of the Atomic Scientists, 61(5):40–49, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/5/40.

Moore:2005:BSC

[Moo05b] LeRoy Moore. The bait-and-switch cleanup. Bulletin of the Atomic Scientists, 61(1):50–57, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/50.

Mooney:2007:IA

[Moo07] Chris Mooney. An inconvenient assessment. Bulletin of the Atomic Scientists, 63(6):40–47, November/December 2007. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/63/6/40.

Moore:2014:CWM

[Moo14] George M. Moore. Out of control: Why mandatory international reporting is needed for radioactive sources and materials. Bulletin of the Atomic Scientists, 70(6):63–72, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moore:2015:CLE

[Moo15] George M. Moore. Could low-enriched uranium be used in naval reactors? Don’t ask the Navy. Bulletin of the Atomic Scientists, 71(3):67–75, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mooney:2018:WCC

[Moo18] Pat Mooney. What’s cooking for climate change? Techno-fixing dinner for 10 billion. Bulletin of the Atomic Scientists, 74(6): 390–396, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1533214.

Moon:2025:GIL

[Moo25] Twila Moon. Greenland ice loss cannot be stopped — but it can and must be slowed. Bulletin of the Atomic Scientists, 81 (2):88–91, 2025. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2025.2464429.

Morrison:1946:LD

[Mor46] Philip Morrison. The Laboratory demobilizes. Bulletin of the Atomic Scientists, 2(9–10):5–6, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1947:ASG

[Mor47a] Philip Morrison. Alsos: The story of German science. Bulletin of the Atomic Scientists, 3(12):354, 365, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [vL47] and reply [Mor48a].

Morrison:1947:RRS

[Mor47b] Philip Morrison. Replies to Ridenour: Science should be kept free. Bulletin of the Atomic Scientists, 3(8):224, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1948:RDL

[Mor48a] Philip Morrison. A reply to Dr. von Laue. Bulletin of the Atomic Scientists, 4(4):104, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab48d, vL48c].

Morse:1948:FTS

[Mor48b] Philip M. Morse. Freedom of thought in science. Bulletin of the Atomic Scientists, 4(11):335, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1949:BAI

[Mor49] Philip Morrison. Blackett’s analysis of the issues. Bulletin of the Atomic Scientists, 5(2):37–40, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgenthau:1950:CUS

[Mor50a] Hans J. Morgenthau. The conquest of the United States by Germany. Bulletin of the Atomic Scientists, 6(1):21–26, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgenthau:1950:HBA

[Mor50b] Hans J. Morgenthau. The H-bomb and after. Bulletin of the Atomic Scientists, 6(3):76–79, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano50m].

Morgenthau:1950:NR

[Mor50c] Hans J. Morgenthau. On negotiating with the Russians. Bulletin of the Atomic Scientists, 6(5):143–148, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1951:ISE

[Mor51] Philip Morrison. II: Science is essentially social. Bulletin of the Atomic Scientists, 7(12):373–375, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mori:1952:LH

[Mor52] Yoshimara Mori. A letter from Hiroshima. Bulletin of the Atomic Scientists, 8(2):53, 64, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moran:1953:CCS

[Mor53a] Lawrence Moran. Catholicism, Communism, and science. Bulletin of the Atomic Scientists, 9(2):49–53, 64, March 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morse:1953:WFP

[Mor53b] Wayne D. Morse. Wanted — fair procedures. Bulletin of the Atomic Scientists, 9(6):211–214, 232, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgenthau:1954:PMS

[Mor54] Hans J. Morgenthau. The political and military strategy of the United States. Bulletin of the Atomic Scientists, 10(8):323–327, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgenthau:1955:ILS

[Mor55] Hans J. Morgenthau. The impact of the loyalty–security measures on the State Department. Bulletin of the Atomic Scientists, 11(4): 134–140, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Morgenthau:1956:AWR

[Mor56] Hans J. Morgenthau. Has atomic war really become impossible? Bulletin of the Atomic Scientists, 12(1):7–9, January 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moraczewski:1957:LEC

[Mor57a] Albert Moraczewski, O.P. Letter to the Editor: The Catholic scientist. Bulletin of the Atomic Scientists, 13(7):273, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moraczewski:1957:PVS

[Mor57b] Father Albert Moraczewski, O.P. Part VII: Science and religion: The contribution of science to religion. Bulletin of the Atomic Scientists, 13(1):27–31, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sac57].

Moraczewski:1959:BBR

[Mor59a] Father Albert Moraczewski, O.P., Ph.D. Books: Religion without Revelation, by Julian Huxley. Bulletin of the Atomic Scientists, 15 (7):315–316, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Cla59, Vis59].

Morgan:1959:HER

[Mor59b] Karl Z. Morgan. Human exposure to radiation. Bulletin of the Atomic Scientists, 15(9):383–389, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morse:1959:MA

[Mor59c] Marston Morse. Mathematics and the arts. Bulletin of the Atomic Scientists, 15(2):55–59, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1961:BBH

[Mor61a] Philip Morrison. Books: No High Ground, by Fletcher Knebel. Bulletin of the Atomic Scientists, 17(1):36–37, January 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morton:1961:LER

[Mor61b] Louis Morton. Letter to the Editor: For the record. Bulletin of the Atomic Scientists, 17(10):429, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moravcsik:1962:LEF

[Mor62a] Michael J. Moravcsik. Letter to the Editor: Further thoughts on shelters. Bulletin of the Atomic Scientists, 18(7):35–36, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys62].

Morgenthau:1962:BBL

[Mor62b] Hans J. Morgenthau. Books: The Legacy of Hiroshima, by Edward Teller with Allen Brown. Bulletin of the Atomic Scientists, 18(6):34–36, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Asi62b].

Morgenthau:1962:DNA

[Mor62c] Hans J. Morgenthau. Decisionmaking in the Nuclear Age. Bulletin of the Atomic Scientists, 18(10):7–8, December 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morton:1962:BBC

[Mor62d] Louis Morton. Books: The Common Defense: Strategic Programs in National Politics, by Samuel P. Huntington; Changing Patterns of Military Politics, edited by Samuel P. Huntington; Organizing for Defense: The American Military Establishment in the Twentieth Century, by Paul Y. Hammond. Bulletin of the Atomic Scientists, 18(5):36–37, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgenthau:1965:VCC

[Mor65] Hans J. Morgenthau. The Vietnam crisis and China. Bulletin of the Atomic Scientists, 21(6):27, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moravcsik:1966:LES

[Mor66a] Michael J. Moravcsik. Letter to the Editor: Scientists in politics. Bulletin of the Atomic Scientists, 22(3):28, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ame66].

Moravcsik:1966:SP

[Mor66b] Michael J. Moravcsik. Scientists in politics — and out. Bulletin of the Atomic Scientists, 22(1):32–34, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgenthau:1967:NFP

[Mor67] Hans J. Morgenthau. A new foreign policy for the United States: Basic issues. Bulletin of the Atomic Scientists, 23(1):7–11, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moravcsik:1970:CNL

[Mor70a] Michael J. Moravcsik. Communications: National laboratories. Bulletin of the Atomic Scientists, 26(8):48, October 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor70b].

Moravcsik:1970:RNL

[Mor70b] Michael J. Moravcsik. Reflections on national laboratories. Bulletin of the Atomic Scientists, 26(2):11–15, February 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addenda [Duf70, Mor70a].

Moravcsik:1971:BDP

[Mor71] Michael J. Moravcsik. On brain drain in the Philippines. Bulletin of the Atomic Scientists, 27(2):36, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mur70].

Moravcsik:1973:TSC

[Mor73] Michael J. Moravcsik. The transmission of a scientific civilization. Bulletin of the Atomic Scientists, 29(3):25–28, March 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pas74].

Morse:1977:TAE

[Mor77] Philip M. Morse. A team approach to energy planning. Bulletin of the Atomic Scientists, 33(3):68–69, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morgan:1978:CLL

[Mor78] Karl Z. Morgan. Cancer and low level ionizing radiation. Bulletin of the Atomic Scientists, 34(7):30–41, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano78j] and comments [CMR79, Coh79].

Morgan:1979:BRB

[Mor79] Robert P. Morgan. Book review: U.S. Science and Technology for Development: A Contribution to the 1979 U.N. Conference, by Joel Bernstein. Bulletin of the Atomic Scientists, 35(6):55–57, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moravcsik:1980:CPN

[Mor80] Michael J. Moravcsik. Commentary: Preparing the new pioneers. Bulletin of the Atomic Scientists, 36(10):56–57, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morell:1982:BRB

[Mor82a] David Morell. Book review: Powerline: The First Battle of America’s Energy War, by Barry M. Casper and Paul David Wellstone. Bulletin of the Atomic Scientists, 38(7):40–41, August/ September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1982:NP

[Mor82b] Philip Morrison. The nuclear plateau. Bulletin of the Atomic Scientists, 38(10):19, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [JM83].

Morgan:1983:SIS

[Mor83a] Robert P. Morgan. Survival is the issue: Sharing science and technology. Bulletin of the Atomic Scientists, 39(5):23–27, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1983:BRB

[Mor83b] David C. Morrison. Book review: The Truth About the Neutron Bomb: The Inventor of the Bomb Speaks Out, by Sam Cohen. Bulletin of the Atomic Scientists, 39(8):36, 38, October 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1983:WTP

[Mor83c] Philip Morrison. The weapons tutorial — Part 1: The spiral of peril: a narrative of the nuclear arms race. Bulletin of the Atomic Scientists, 39(1):10–17, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Parts 2 [Tsi83i], 3 [Mar83b], 4 [Hol83c], 5 [Sar83], and 6 [Rom83].

Morrison:1984:BRB

[Mor84a] David C. Morrison. Book review: The Command and Control of Nuclear Forces, by Paul Bracken. Bulletin of the Atomic Scientists, 40(4):35–36, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1984:IV

[Mor84b] David C. Morrison. ICBM vulnerability. Bulletin of the Atomic Scientists, 40(9):22–29, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BR85].

Morrison:1984:WTP

[Mor84c] David C. Morrison. The weapons tutorial: [part 8:] air-breathing nuclear delivery systems. Bulletin of the Atomic Scientists, 40 (2):32–39, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morris:1985:BRB

[Mor85a] Valarie Ziegler Morris. Book reviews: Rebels Against War: The American Peace Movement, 1933–1983, by Lawrence S. Wittner; The Politics of the Nuclear Freeze, by Adam M. Garfinkle. Bulletin of the Atomic Scientists, 41(8):57–59, September 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1985:EDW

[Mor85b] David C. Morrison. Energy Department’s weapons conglomerate. Bulletin of the Atomic Scientists, 41(4):32–37, April 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrison:1985:NAJ

[Mor85c] David C. Morrison. Nuclear allergy: Japanese principles, U.S. policies. Bulletin of the Atomic Scientists, 41(6):22–24, June/ July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morel:1987:ASS

[Mor87a] Benoit Morel. ATBM — a solution in search of a problem. Bulletin of the Atomic Scientists, 43(4):39–41, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morland:1987:LAC

[Mor87b] Howard Morland. Letter: Arms control matters. Bulletin of the Atomic Scientists, 43(2):61, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mcc86c].

Morrison:1987:BRB

[Mor87c] Roy Morrison. Book review: America Can Win: The Case for Military Reform, by Gary Hart with William S. Lind. Bulletin of the Atomic Scientists, 43(8):49–51, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morrow:1988:ME

[Mor88] Thomas H. Morrow. Morrow’s evidence. Bulletin of the Atomic Scientists, 44(8):16, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morland:2000:WLP

[Mor00] Howard Morland. What’s left to protect? Bulletin of the Atomic Scientists, 56(6):51–55, November/December 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/6/51.

Moran:2009:APN

[Mor09] Barbara Moran. The aftermath of the Palomares nuclear accident. Bulletin of the Atomic Scientists, 65(3):48–54, May 2009.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/3/48.

Morozov:2012:AWC

[Mor12] Yury Morozov. Arctic 2030: What are the consequences of climate change?: The Russian response. Bulletin of the Atomic Scientists, 68(4):22–27, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/4/22.

Moreira:2014:BEI

[Mor14] José R. Moreira. Biomass energy and the implications for climate and food: The Brazilian response. Bulletin of the Atomic Scientists, 70(1):12–15, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/70/1/12.full.pdf+html.

Moshinsky:1952:TGN

[Mos52] Marcus Moshinsky. The treatment of good neighbors. Bulletin of the Atomic Scientists, 8(7):251, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Moss:1961:CS

[Mos61] John E. Moss. The crisis of secrecy. Bulletin of the Atomic Scientists, 17(1):8–11, 35, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1965:LEM

[Mos65] Norman Moss. Letter to the Editor: More on Levine. Bulletin of the Atomic Scientists, 21(8):28–29, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev64].

Moss:1967:SSM

[Mos67] Laurence I. Moss. In searched of a subsidy machine: Or, why the Grand Canyon must be dammed. Bulletin of the Atomic Scientists, 23(6):25–30, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moscone:1979:VSN

[Mos79] George R. Moscone. A voice is silenced: November 27, 1978. Bulletin of the Atomic Scientists, 35(1):64, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1980:CLG

[Mos80] Norman Moss. Commentary from London: A global code for nuclear fuel. Bulletin of the Atomic Scientists, 36(2):8–11, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1981:BRBb

[Mos81a] Norman Moss. Book review: Britain and Nuclear Weapons, by Lawrence Freedman. Bulletin of the Atomic Scientists, 37(6):45– 46, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1981:BRBa

[Mos81b] Norman Moss. Book review: Nuclear Energy and Nuclear Proliferation: Japanese and American Views, by Ryukichi Imai and Henry S. Rowan. Bulletin of the Atomic Scientists, 37(4):45–46, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moshinsky:1984:WUS

[Mos84] Marcos Moshinsky. On: ‘Why the United States is in Central America’. Bulletin of the Atomic Scientists, 40(2):58–59, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ken83].

Moss:1985:LTH

[Mos85] Norman Moss. Letter: Truman and the H-Bomb. Bulletin of the Atomic Scientists, 41(2):53, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1988:VIB

[Mos88] Norman Moss. Vanunu, Israel’s bombs and US aid. Bulletin of the Atomic Scientists, 44(4):7–8, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [FP89].

Moss:1989:RLU

[Mos89] Norman Moss. Reports from London: Unilateral disarmament — Labor’s lost love. Bulletin of the Atomic Scientists, 45(8):9– 11, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gap90].

Moss:1990:BRH

[Mos90] Norman Moss. Book review: High spies: America’s Secret Eyes in Space, by Jeffrey T. Richelson. Bulletin of the Atomic Scientists, 46(9):43, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1991:BRM

[Mos91] Norman Moss. Book review: Mossad mayhem: Every Spy a Prince: The Complete History of Israel’s Intelligence Community, by Dan Raviv and Yossi Melman. Bulletin of the Atomic Scientists, 47(1):44, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mosher:1992:BRW

[Mos92] Lawrence Mosher. Book review: When fences fail: Trinity’s Children: Living Along America’s Nuclear Highway, by Tad Bartimus and Scott McCartney. Bulletin of the Atomic Scientists, 48(9):42– 43, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1993:RBT

[Mos93a] Norman Moss. Report: Britain: THORP flap. Bulletin of the Atomic Scientists, 49(2):8–9, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moss:1993:RSS

[Mos93b] Norman Moss. Report: Spies: “Sonya” explains. Bulletin of the Atomic Scientists, 49(6):9–11, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1947:ICC

[Mot47] N. F. Mott. International control: The choice before the country. Bulletin of the Atomic Scientists, 3(11):315–320, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1948:AAH

[Mot48] Nevill F. Mott. An atomic alliance holds no security for Britain. Bulletin of the Atomic Scientists, 4(12):364–365, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fri48a].

Mott:1949:CAW

[Mot49] N. F. Mott. Can atomic weapons keep the peace? Bulletin of the Atomic Scientists, 5(1):11–12, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1951:IWS

[Mot51] N. F. Mott. III. Working for a society where science can thrive. Bulletin of the Atomic Scientists, 7(12):375–376, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mottet:1964:LEM

[Mot64] Arthur L. Mottet, Jr. Letter to the Editor: Materialism. Bulletin of the Atomic Scientists, 20(4):34, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1968:BBB

[Mot68] Nevill F. Mott. Books: Reflections on Big Science, by Alvin M. Weinberg. Bulletin of the Atomic Scientists, 24(5):38–39, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1982:EON

[Mot82] Sir Nevill Mott. European opinion and NATO policy. Bulletin of the Atomic Scientists, 38(6):4–5, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1983:CCU

[Mot83] Sir Nevill Mott. Commentary: The controversy in the United Kingdom: No first use. Bulletin of the Atomic Scientists, 39(7): 53–54, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mott:1985:LFI

[Mot85] Nevill Mott. Letter: Falklands insensitivity. Bulletin of the Atomic Scientists, 41(2):56, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bal84b].

Motyl:1989:SRI

[Mot89] Alexander J. Motyl. Soviet republics: Identity crisis in the Soviet west. Bulletin of the Atomic Scientists, 45(2):21–24, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mouat:1973:CGL

[Mou73] Malcolm P. Mouat. Communications: Global legal disputes. Bulletin of the Atomic Scientists, 29(3):2, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mount:2017:ANM

[Mou17] Adam Mount. Adapting nuclear modernization to the new administration. Bulletin of the Atomic Scientists, 73(3):167–172, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1315036.

Mowlam:1983:EPG

[Mow83] Marjorie Mowlam. From Europe: Peace groups and politics. Bulletin of the Atomic Scientists, 39(9):28–32, November 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mowlana:1995:CP

[Mow95a] Hamid Mowlana. The communications paradox. Bulletin of the Atomic Scientists, 51(4):40–46, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mowlana:1995:IE

[Mow95b] Hamid Mowlana. The Internet elite. Bulletin of the Atomic Scientists, 51(4):44–45, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mackintosh:1978:CWC

[MP78] A. R. Mackintosh and Walter C. Patterson. Commentary: Windscale critique. Bulletin of the Atomic Scientists, 34(10):57–58, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McCahon:1977:NSP

[MR77] Jim McCahon and Murray Robertson. The nuclear South Pacific. Bulletin of the Atomic Scientists, 33(4):26–27, April 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marples:1990:RCI

[MR90] David R. Marples and Yuri Risovanny. Revelations of a Chernobyl insider. Bulletin of the Atomic Scientists, 46(10):16–21, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mian:2007:TF

[MR07] Zia Mian and M. V. Ramana. Triumph of fear. Bulletin of the Atomic Scientists, 63(4):34, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/4/34.

Mian:2009:NSE

[MR09] Zia Mian and M. V. Ramana. A nuclear and sustainable energy reading list. Bulletin of the Atomic Scientists, 65(6):78–85, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/65/
6/78.

Makhijani:2021:CSM

[MR21] Arjun Makhijani and M. V. Ramana. Can small modular reactors help mitigate climate change? Bulletin of the Atomic Scientists, 77(4):207–214, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1941600.

Mostafa:2016:CAF

[MRH16] Meraz Mostafa, M. Feisal Rahman, and Saleemul Huq. Climate adaptation funding: Getting the money to those who need it. Bulletin of the Atomic Scientists, 72(6):396–401, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Martin:1961:RCT

[MS61] Paul S. Martin and Cornelius Steelink. Readers’ comments: The Titanizing of Tucson. Bulletin of the Atomic Scientists, 17(4):

167–168, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milstein:1977:UMR

[MS77a] M. A. Milstein and L. S. Semejko. U.S. military R&D through Soviet eyes. Bulletin of the Atomic Scientists, 33(2):32–38, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morehouse:1977:LRS

[MS77b] Ware Morehouse and Jon Sigurdson. Linkages of R&D systems to contemporary societies. Bulletin of the Atomic Scientists, 33 (10):26–27, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Morehouse:1977:STP

[MS77c] Ware Morehouse and Jon Sigurdson. Science, technology and poverty: issues underlying the 1979 U.N. Conference on Science and Technology for Development. Bulletin of the Atomic Scientists, 33(10):21–28, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McKim:1985:LID

[MS85] Robert McKim and Robert A. Seeley. Letter: Immorality of deterrence. Bulletin of the Atomic Scientists, 41(5):52–53, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Per85].

Marsh:2006:BI

[MS06] Gerald E. Marsh and George S. Stanford. Batteries included. Bulletin of the Atomic Scientists, 62(6):19–20, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/6/19.

Marsh:1980:CFB

[MSDP80] G. E. Marsh, G. S. Stanford, A. De Volpi, and T. A. Postol. Commentary: Fallout from the Progressive H-bomb. Bulletin of the Atomic Scientists, 36(4):64–65, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79i].

McKinney:2020:WAB

[MSW20] Katherine E. McKinney, Scott D. Sagan, and Allen S. Weiner. Why the atomic bombing of Hiroshima would be illegal today. Bulletin of the Atomic Scientists, 76(4):157–165, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1778344.

Marks:1955:CIA

[MT55] Herbert S. Marks and George F. Trowbridge. Control of information and the Atomic Energy Act of 1954. Bulletin of the Atomic Scientists, 11(4):128–130, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Morrison:1996:LTT

[MT96] Philip Morrison and Kosta Tsipis. Letter: Telling time. Bulletin of the Atomic Scientists, 52(3):60, May/June 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark96d].

Morrison:2003:RN

[MT03] Philip Morrison and Kosta Tsipis. Rightful names. Bulletin of the Atomic Scientists, 59(3):77, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/3/77.

Mtimkulu:1996:RNA

[Mti96a] Bereng Mtimkulu. Report: NWFZ: Africa bans the bomb. Bulletin of the Atomic Scientists, 52(4):11, July/August 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mtimkulu:1996:RSA

[Mti96b] Bereng Mtimkulu. Report: South Africa: Reluctant peacekeeper. Bulletin of the Atomic Scientists, 52(1):9–10, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Marshak:1946:DCI

[MTK46] J. Marshak, E. Teller, and L. R. Klein. Dispersal of cities and industries. Bulletin of the Atomic Scientists, 1(9):13–15, 20, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mueller:1969:IMP

[Mue69] George E. Mueller. The investigation of the Moon: The 1970– 71 plan. Bulletin of the Atomic Scientists, 25(7):67, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mueller:1977:LSF

[Mue77] Ruth Mueller. Letter: Sheets and fireballs. Bulletin of the Atomic Scientists, 33(9):2, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cla77].

Mueller:2001:MDG

[Mue01] Harald Mueller. Missile defense: Germany hopes it will go away. Bulletin of the Atomic Scientists, 57(6):31–33, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/6/31.

Mueller:2007:FNN

[Mue07a] John Mueller. Fear not: Notes from a naysayer. Bulletin of the Atomic Scientists, 63(2):??, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/2/30.

Mueller:2007:FFI

[Mue07b] John Mueller. Fearing fear itself. Bulletin of the Atomic Scientists, 63(6):55–56, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/6/55.

Mukherjee:1993:LIN

[Muk93] Sadhan Mukherjee. Letter: India and the NPT. Bulletin of the Atomic Scientists, 49(1):54–55, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AH92d].

Muller:1947:CGT

[Mul47] H. J. Muller. Changing genes: Their effects on evolution. Bulletin of the Atomic Scientists, 3(9):267–272, 274, September 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1948:CGU

[Mul48] H. J. Muller. The crushing of genetics in the USSR. Bulletin of the Atomic Scientists, 4(12):369–371, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1952:WSC

[Mul52] H. J. Muller. Will science continue? Bulletin of the Atomic Scientists, 8(9):301–307, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mulgrue:1953:WBA

[Mul53] George Mulgrue. The World’s battle against hunger. Bulletin of the Atomic Scientists, 9(5):197–198, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1955:HRC

[Mul55a] H. J. Muller. How radiation changes the genetic constitution. Bulletin of the Atomic Scientists, 11(9):329–338, 352, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1955:GDP

[Mul55b] J. J. Muller. The genetic damage produced by radiation. Bulletin of the Atomic Scientists, 11(6):210–212, 230, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1959:SH

[Mul59] H. J. Muller. Science for humanity. Bulletin of the Atomic Scientists, 15(4):146–150, 176, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1960:MF

[Mul60] H. J. Muller. The meaning of freedom. Bulletin of the Atomic Scientists, 16(8):311–316, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Har60b].

Muller:1964:PLS

[Mul64a] H. J. Muller. Perspectives for the life sciences. Bulletin of the Atomic Scientists, 20(1):3–7, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [Elw64].

Muller:1964:RR

[Mul64b] Steven Muller. A reawakening on the Rhine. Bulletin of the Atomic Scientists, 20(7):2–6, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tuc64].

Muller:1965:BBT

[Mul65] Steven Muller. Books: The Troubled Partnership: A Re-Appraisal of the Atlantic Alliance, by Henry A. Kissinger; Alternative to Partition: For a Broader Conception of America’s Role in Europe, by Zbigniew Brzezinski; After Twenty Years: Alternatives to the Cold War in Europe, by Richard J. Barnet and Marcus G. Raskin. Bulletin of the Atomic Scientists, 21(9):26–28, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1966:BBG

[Mul66] Steven Muller. Books: The Grand Design: A European Solution to German Reunification, by Franz Josef Strauss; The Western Alliance: Its Status and Prospects, by Edgar S. Furniss. Bulletin of the Atomic Scientists, 22(5):33–35, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1967:BST

[Mul67] Herman J. Muller. Biologists’ statement on teaching evolution. Bulletin of the Atomic Scientists, 23(2):39–40, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter signed by 177 scientists.

Muller:1968:WGC

[Mul68] H. J. Muller. What genetic course will man steer? Bulletin of the Atomic Scientists, 24(3):6–12, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Published posthumously, but first presented to the 1966 International Congress on Human Genetics.

Muller:1969:ABG

[Mul69] Steven Muller. Anxiety in Bonn: German fears after Czechoslovakia. Bulletin of the Atomic Scientists, 25(3):13–15, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1971:CS

[Mul71] Rene J. Muller. Communications: A scenario for 2200. Bulletin of the Atomic Scientists, 27(9):2–3, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1972:BBO

[Mul72] Steven Muller. Books: Germany in Our Time, by Alfred Grosser. Bulletin of the Atomic Scientists, 28(4):46–47, April 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1973:RRR

[Mul73a] Herbert J. Muller. Reflections on re-reading Darwin. Bulletin of the Atomic Scientists, 29(2):5–8, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1973:BBB

[Mul73b] Rene J. Muller. Books: In Bluebeard’s Castle: Some Notes Toward the Redefinition of Culture, by George Steiner. Bulletin of the Atomic Scientists, 29(3):53–56, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mulliken:1974:LR

[Mul74a] Robert S. Mulliken. Letters: Response. Bulletin of the Atomic Scientists, 30(9):3, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul74b, BCW74].

Mulliken:1974:TZN

[Mul74b] Robert S. Mulliken. Through ZPG to NPG. Bulletin of the Atomic Scientists, 30(1):9–11, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bra74, Tyn75], correction [BCW74], and response [Mul74a].

Mulliken:1977:LSS

[Mul77] Robert S. Mulliken. Letter: Science and sense. Bulletin of the Atomic Scientists, 33(2):4, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei76b].

Mulliken:1980:CCW

[Mul80] Robert S. Mulliken. Commentary: The cold war. Bulletin of the Atomic Scientists, 36(8):8–10, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1985:SUN

[Mül85] Harald Müller. Superpowers’ unfulfilled NPT promise. Bulletin of the Atomic Scientists, 41(8):18–20, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1989:LBC

[Mül89] Harald Müller. Letter: Bundestag committee worked. Bulletin of the Atomic Scientists, 45(7):43, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha89b].

Muller:1990:NRW

[Mül90a] Harald Müller. NPT review: Western Europe needs treaty. Bulletin of the Atomic Scientists, 46(6):28–29, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1990:PNN

[Mül90b] Harald Müller. Perspective: No need to fear united Germany. Bulletin of the Atomic Scientists, 46(3):14–15, April 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muller:1993:ELB

[Mül93] Harald Müller. Europe’s leaky borders. Bulletin of the Atomic Scientists, 49(5):27–29, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muldoon:1995:WHH

[Mul95] James P. Muldoon. What happened to humanitarian intervention? Bulletin of the Atomic Scientists, 51(2):60–61, March/ April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mumford:1954:ASA

[Mum54] Lewis Mumford. Anticipations and social adjustments in science. Bulletin of the Atomic Scientists, 10(2):34–36, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Munger:1954:BBS

[Mun54] Edwin S. Munger. Books: Africa — A Study in Tropical Development, by L. Dudley Stamp. Bulletin of the Atomic Scientists, 10(3):99, 102, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Munk:1965:LER

[Mun65] Arthur W. Munk. Letter to the Editor: The road to peace and freedom. Bulletin of the Atomic Scientists, 21(8):30–31, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Munday:1967:CU

[Mun67a] John C. Munday, Jr. Comments: On the UFOs. Bulletin of the Atomic Scientists, 23(10):40–41, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar67a, Sag67].

Munk:1967:CP

[Mun67b] Arthur Munk. Comments: Priorities. Bulletin of the Atomic Scientists, 23(2):27, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murphy:1951:GEP

[Mur51] Walter J. Murphy. Guest editorial: Postponement, not exemption. Bulletin of the Atomic Scientists, 7(5):130, May 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murray:1954:CO

[Mur54a] Thomas E. Murray. Concurring opinion. Bulletin of the Atomic Scientists, 10(7):277–279, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murray:1954:CC

[Mur54b] Thomas E. Murray. Current comment. Bulletin of the Atomic Scientists, 10(1):31, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murray:1954:DLA

[Mur54c] Thomas E. Murray. Don’t leave atomic energy to the experts. Bulletin of the Atomic Scientists, 10(2):48–50, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murray:1955:ODL

[Mur55a] Thomas E. Murray. One definition of loyalty. Bulletin of the Atomic Scientists, 11(4):125, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Murrow:1955:TDL

[Mur55b] Edward R. Murrow. Is there a dividing line? Bulletin of the Atomic Scientists, 11(5):192, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Taken from the author’s broadcast on CBS, 10 March 1955.

Murray:1958:GD

[Mur58] Thomas E. Murray. Gordon Dean 1905–1958. Bulletin of the Atomic Scientists, 14(8):335, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murray:1960:EET

[Mur60] Thomas E. Murray. Ending an era of terror. Bulletin of the Atomic Scientists, 16(5):162–166, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murcier:1968:RBS

[Mur68] Alain Murcier. Reports: Brains for sale. Bulletin of the Atomic Scientists, 24(3):38–46, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muriel:1970:BDP

[Mur70] Amador Muriel. Brain drain in the Philippines: a case study. Bulletin of the Atomic Scientists, 26(7):38–39, September 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Mor71].

Murphy:1972:MCE

[Mur72] Charles H. Murphy. Mainland China’s evolving nuclear deterrent. Bulletin of the Atomic Scientists, 28(1):28–35, January 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murley:1975:LHU

[Mur75] Thomas E. Murley. Letters: Hard to understand. Bulletin of the Atomic Scientists, 31(7):4, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Murphy:1984:FRB

[Mur84] Mary C. Murphy. Film review: A Call for Peace: The Military Budget and You. Bulletin of the Atomic Scientists, 40(3):50, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muskie:1965:MR

[Mus65] Edmund S. Muskie. Muskie requests. Bulletin of the Atomic Scientists, 21(9):25, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muskrat:1970:LA

[Mus70] Joseph C. Muskrat. Landless in Alaska. Bulletin of the Atomic Scientists, 26(3):12–16, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mustard:1975:LFF

[Mus75] John K. Mustard. Letters: Fossil foolishness. Bulletin of the Atomic Scientists, 31(6):3, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol75].

Muskat:1976:LDD

[Mus76] Morris Muskat. Letter: Dilemma of deterrence. Bulletin of the Atomic Scientists, 32(9):3–4, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Musante:1978:LA

[Mus78] Frederick L. Musante, Jr. Letter: In appreciation. Bulletin of the Atomic Scientists, 34(2):7, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Agn77b].

Mustard:1979:LAF

[Mus79] John K. Mustard. Letter: Auto fuels. Bulletin of the Atomic Scientists, 35(3):87, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Muskat:1980:CID

[Mus80] Morris Muskat. Commentary: Illusions of decontrol. Bulletin of the Atomic Scientists, 36(2):59–61, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Music:1982:GN

[Mus82] Robert K. Music. Growing up nuclear. Bulletin of the Atomic Scientists, 38(1):19, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mushtaq:1999:FKD

[Mus99a] Najum Mushtaq. A funny kind of deterrence. Bulletin of the Atomic Scientists, 55(3):6, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/3/6.

Mushtaq:1999:ID

[Mus99b] Najum Mushtaq. Islam distorted. Bulletin of the Atomic Scientists, 55(4):53–57, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/4/53.

Mushtaq:2000:PMJ

[Mus00] Najum Mushtaq. Pakistan: Moderate jihad? Bulletin of the Atomic Scientists, 56(4):15–16, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/4/15.

Mushtaq:2001:PKF

[Mus01] Najum Mushtaq. Pakistan: Khan forced out. Bulletin of the Atomic Scientists, 57(4):13–15, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/4/13.

Mushtaq:2002:LP

[Mus02] Najum Mushtaq. Letter from Pakistan. Bulletin of the Atomic Scientists, 58(2):27–29, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/2/27.

Musil:2004:PP

[Mus04] Robert K. Musil. People power. Bulletin of the Atomic Scientists, 60(1):69–70, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/1/69.

Mushtaq:2007:DC

[Mus07] Najum Mushtaq. Divide and conquer. Bulletin of the Atomic Scientists, 63(1):17–19, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/17.

Maerli:2004:ECY

[MV04] Morten Bremer Mærli and Lars Van Dassen. Europe, carry your weight. Bulletin of the Atomic Scientists, 60(6):19–21, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/6/19.

Maurer:2013:ISG

[MvE13] Stephen M. Maurer and Sebastian von Engelhardt. Industry selfgovernance: a new way to manage dangerous technologies. Bulletin of the Atomic Scientists, 69(3):53–62, May/June 2013. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/69/3/53.

Miller:1997:PLR

[MvH97] Marvin Miller and Frank von Hippel. Perspective: Let’s reprocess the MOX plan. Bulletin of the Atomic Scientists, 53(4):15–17, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Macfarlane:2001:PDT

[MvHKN01] Allison Macfarlane, Frank von Hippel, Jungmin Kang, and Robert Nelson. Plutonium disposal, the third way. Bulletin of the Atomic Scientists, 57(3):53–57, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/53.

Morrison:1947:HWL

[MW47] Philip Morrison and Robert R. Wilson. Half a world … and none: Partial world government criticized. Bulletin of the Atomic Scientists, 3(7):181–182, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Montgomery:1972:CEC

[MW72] D. J. Montgomery and Stan Wiggins. Communications: Ehrlich and Commoner. Bulletin of the Atomic Scientists, 28(7):2, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [EH72a, Com72].

Murphey:1975:LAN

[MW75] Walter Murphey and Eugene P. Wigner. Letters: Armed neutrality. Bulletin of the Atomic Scientists, 31(6):2, June 1975. CODEN BASIAP. ISSN 0096-3402.

Meek:1977:LNM

[MW77] Daniel Meek and Alvin M. Weinberg. Letter: Nuclear misconceptions. Bulletin of the Atomic Scientists, 33(6):3–5, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei77a].

Meerburg:1981:CCD

[MW81] Arend J. Meerburg and Arthur H. Westing. Commentary: Crop destruction. Bulletin of the Atomic Scientists, 37(8):61, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wes81].

Meader:1983:CER

[MW83] Norman M. Meader and Alvin Weinberg. Commentary: The entropists revisited. Bulletin of the Atomic Scientists, 39(3):61– 62, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei82a].

Milhollin:1991:RNP

[MW91] Gary Milhollin and Gerard White. Report: Nuclear proliferation: Tug of war over high-tech experts. Bulletin of the Atomic Scientists, 47(2):7–8, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Milne:1998:AND

[MW98] Tom Milne and Henrietta Wilson. Aldermaston and nuclear disarmament. Bulletin of the Atomic Scientists, 54(6):24–25, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maddox:1967:CPU

[MY67] John Maddox and Elizabeth Young. Comments: Prestige unimportant. Bulletin of the Atomic Scientists, 23(5):63–64, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1975:LNL

[Mye75] Edward A. Myers, Jr. Letters: A new low. Bulletin of the Atomic Scientists, 31(1):2, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch74b, Day75l].

Myers:1982:REL

[Mye82a] Dave Myers. Rabinowitch Essay: The last Europe. Bulletin of the Atomic Scientists, 38(3):22–25, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1982:RA

[Mye82b] Norman Myers. Room in the ark? Bulletin of the Atomic Scientists, 38(9):44–48, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1986:SWS

[Mye86] Ware Myers. The Star Wars software debate. Bulletin of the Atomic Scientists, 42(2):31–36, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wal86].

Myers:1987:BRB

[Mye87a] Norman Myers. Book reviews: Global Resources and International Conflict: Environmental Factors in Strategic Policy and Action, edited by Arthur H. Westing; Population Growth in Latin America and U.S. National Security, edited by John Saunders. Bulletin of the Atomic Scientists, 43(4):52–53, May 1, 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1987:LES

[Mye87b] Norman Myers. Linking environment and security. Bulletin of the Atomic Scientists, 43(5):46–47, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1988:BRB

[Mye88] Norman Myers. Book review: World Encyclopedia of Peace, edited by Ervin Laszlo and Jong Youl Yoo. Bulletin of the Atomic Scientists, 44(3):54–55, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1989:BRB

[Mye89] Norman Myers. Book review: The Politics of Scarcity: Water in the Middle East, edited by Joyce R. Starr and Daniel C. Stoll. Bulletin of the Atomic Scientists, 45(5):42–43, June 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1990:LDF

[Mye90] R. Thomas Myers. Letter: Disarmament by fixed fraction, free choice. Bulletin of the Atomic Scientists, 46(2):46–47, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kra89].

Myers:1991:LMP

[Mye91] R. Thomas Myers. Letter: Manhattan Project for renewable energy. Bulletin of the Atomic Scientists, 47(9):46, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [AG91].

Myers:1992:RUC

[Mye92] Nancy Myers. Report: Ukraine: Coping with Chernobyl. Bulletin of the Atomic Scientists, 48(7):8–9, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1993:ENR

[Mye93a] Nancy Myers. Editor’s note: Rediscovering Russia. Bulletin of the Atomic Scientists, 49(1):2, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1993:BRF

[Mye93b] Norman Myers. Book review: Final warning: China’s Environmental Crisis, by Vaclav Smil. Bulletin of the Atomic Scientists, 49(10):50–51, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:1995:WTI

[Mye95] Nancy J. Myers. What time is it, really? Bulletin of the Atomic Scientists, 51(5):8–9, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myers:2002:WM

[Mye02] Norman Myers. Water and more. Bulletin of the Atomic Scientists, 58(5):69–70, September/October 2002. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/5/69.

Mynett:2011:LCC

[Myn11] Arthur Mynett. Lessons of climate change, stories of solutions: The Netherlands: Innovative technology. Bulletin of the Atomic Scientists, 67(1):51–55, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/67/1/51.

Myrdal:1966:RCS

[Myr66] Jan Myrdal. The reshaping of Chinese society. Bulletin of the Atomic Scientists, 22(6):76–79, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1968:TLP

[Myr68] Gunnar Myrdal. Too late to plan? Bulletin of the Atomic Scientists, 24(1):5–9, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1969:SSS

[Myr69] Gunnar K. Myrdal. The soft states of South Asia: The civil servant problem. Bulletin of the Atomic Scientists, 25(4):7–10, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1971:DCR

[Myr71] Gunnar Myrdal. Are the developing countries really developing? Bulletin of the Atomic Scientists, 27(1):5–8, January 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1972:GMC

[Myr72] Gunnar Myrdal. Gunnar Myrdal comments on America’s image, black rebellion, limits to growth and population control. Bulletin of the Atomic Scientists, 28(9):5–7, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lei73].

Myrdal:1973:HSS

[Myr73] Gunnar Myrdal. How scientific are the social sciences? Bulletin of the Atomic Scientists, 29(1):31–37, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1974:EFW

[Myr74] Gunnar Myrdal. Editorial: First we must change society. Bulletin of the Atomic Scientists, 30(6):36–37, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1975:PNE

[Myr75] Alvy Myrdal. ‘Peaceful’ nuclear explosions: Nations must seal off this back-door approach to acquiring weapons stockpiles. Bulletin of the Atomic Scientists, 31(5):29–33, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Myrdal:1980:GEE

[Myr80] Alva Myrdal. Guest editorial: Europe as hostage of the superpowers? Bulletin of the Atomic Scientists, 36(4):4–6, April 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Makhijani:1996:UCI

[MZ96] Arjun Makhijani and Hisham Zerriffi. The U.S. can’t have it both ways. Bulletin of the Atomic Scientists, 52(2):36–39, March/ April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Maurer:2007:SB

[MZ07] Stephen M. Maurer and Laurie Zoloth. Synthesizing biosecurity. Bulletin of the Atomic Scientists, 63(6):16–18, November/ December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
6/16.

Makhijani:2001:MTA

[MZM01] Arjun Makhijani, Hisham Zerriffi, and Annie Makhijani. Magical thinking: Another go at transmutation. Bulletin of the Atomic Scientists, 57(2):34–41, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/2/34.

Norris:1987:EPU

[NA87a] Robert S. Norris and William M. Arkin. Explosive power of the U.S. nuclear weapons stockpile. Bulletin of the Atomic Scientists, 43(6):64, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ibs87].

Norris:1987:NCS

[NA87b] Robert S. Norris and William M. Arkin. Nuclear-capable ships and submarines (1987). Bulletin of the Atomic Scientists, 43(7): 63, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNB

[NA87c] Robert S. Norris and William M. Arkin. Nuclear notebook: British, French, and Chinese nuclear weapons. Bulletin of the Atomic Scientists, 43(10):52, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNCc

[NA87d] Robert S. Norris and William M. Arkin. Nuclear notebook: Chewing on Trident; Bird? Plane? Nuclear strike?; Tomahawk warpath; weapons watch. Bulletin of the Atomic Scientists, 43 (10):51, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNCa

[NA87e] Robert S. Norris and William M. Arkin. Nuclear notebook: Cocked pistols in Europe; stealthy, hypersonic cruise missiles? Lance missiles to South Korea; weapons watch. Bulletin of the Atomic Scientists, 43(4):56, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNCb

[NA87f] Robert S. Norris and William M. Arkin. Nuclear notebook: Command, contra, communications; Dungeons and Dragons; surf and turf; countering Star Warski; weapons watch. Bulletin of the Atomic Scientists, 43(9):55, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNL

[NA87g] Robert S. Norris and William M. Arkin. Nuclear notebook: Last of the Titans; Soviet Navy update; choosing warhead designs; weapons watch. Bulletin of the Atomic Scientists, 43(7): 62, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNM

[NA87h] Robert S. Norris and William M. Arkin. Nuclear notebook: Missile mania; nukes of Earle; expensive bargaining chips; weapons watch; U.S. and Soviet nuclear weapons under development. Bulletin of the Atomic Scientists, 43(8):55–56, October 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNN

[NA87i] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear threat to Libya? a new ICBM capability; SIOP revised; new bomber for Europe; son of SRAM; weapons watch. Bulletin of the Atomic Scientists, 43(5):55, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NNT

[NA87j] Robert S. Norris and William M. Arkin. Nuclear notebook: Trouble with Tomahawk?; new medium machine builder; secret test monitors; new Mirages on the horizon. Bulletin of the Atomic Scientists, 43(6):63, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:NWS

[NA87k] Robert S. Norris and William M. Arkin. Nuclear weapons at sea (1987). Bulletin of the Atomic Scientists, 43(7):63, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:USN

[NA87l] Robert S. Norris and William M. Arkin. U.S. and Soviet nuclearcapable aircraft (1987). Bulletin of the Atomic Scientists, 43(9): 56, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNUb

[NA88a] Robert S. Norris and William M. Arkin. Nuclear notebook: 23,400 [U.S. warheads] and counting; black budget blues; souther flank wrapup; weapons watch. Bulletin of the Atomic Scientists, 44(2):55, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNCa

[NA88b] Robert S. Norris and William M. Arkin. Nuclear notebook: Chasing Bears; off the ground, into the air; the lessons of October. Bulletin of the Atomic Scientists, 44(1):55, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNCb

[NA88c] Robert S. Norris and William M. Arkin. Nuclear notebook: Compensating for INF; Soviet naval developments; nuclear diplomacy; weapons watch. Bulletin of the Atomic Scientists, 44(3):59, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNE

[NA88d] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated Soviet nuclear stockpile, July 1988. Bulletin of the Atomic Scientists, 44(6):56, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNK

[NA88e] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests by year, 1945 to December 31, 1987. Bulletin of the Atomic Scientists, 44(2):56, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNMa

[NA88f] Robert S. Norris and William M. Arkin. Nuclear notebook: Maritime strategy submerges; Soviet military lethargy; weapons watch. Bulletin of the Atomic Scientists, 44(8):55, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNMb

[NA88g] Robert S. Norris and William M. Arkin. Nuclear notebook: Mighty bombing run; cruisin’ for a bruisin’; weapons watch. Bulletin of the Atomic Scientists, 44(10):51, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNNb

[NA88h] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear layoffs; paper missiles; sub nationalism; weapons watch.

Bulletin of the Atomic Scientists, 44(5):55, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNNa

[NA88i] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear pursuits. Bulletin of the Atomic Scientists, 44(3):60, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNNc

[NA88j] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear weapons at sea, 1988. Bulletin of the Atomic Scientists, 44(7):64, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNSb

[NA88k] Robert S. Norris and William M. Arkin. Nuclear notebook: Off to a slow START; bombed out; fewer bucks for the bang; hatchet jobs. Bulletin of the Atomic Scientists, 44(6):55, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNPa

[NA88l] Robert S. Norris and William M. Arkin. Nuclear notebook: Pacific provocations; missiles away; weapons watch. Bulletin of the Atomic Scientists, 44(7):63, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNPb

[NA88m] Robert S. Norris and William M. Arkin. Nuclear notebook: Power proliferation; Turkey balks; weapons watch. Bulletin of the Atomic Scientists, 44(9):47, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNR

[NA88n] Robert S. Norris and William M. Arkin. Nuclear notebook: Ronald Reagan’s military budget dreams. Bulletin of the Atomic Scientists, 44(10):52, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNSa

[NA88o] Robert S. Norris and William M. Arkin. Nuclear notebook: Safety second; nuclear shopping list; not making wave; warhead retreads; short-range attack on START. Bulletin of the Atomic Scientists, 44(4):55, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNUf

[NA88p] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. and Soviet nuclear-capable aircraft (1988). Bulletin of the Atomic Scientists, 44(9):48, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNUe

[NA88q] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. and Soviet nuclear weapons under development, 1988. Bulletin of the Atomic Scientists, 44(8):56, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNUc

[NA88r] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. and Soviet strategic nuclear forces, 1972–1987. Bulletin of the Atomic Scientists, 44(4):56, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNUa

[NA88s] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. and Soviet strategic nuclear forces, end of 1987. Bulletin of the Atomic Scientists, 44(1):56, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1988:NNUd

[NA88t] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile (June 1988). Bulletin of the Atomic Scientists, 44(5):56, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNMb

[NA89a] Robert S. Norris and William M. Arkin. Nuclear notebook: 10megaton quake; Rocky [Flats] road; not superplane?; ELF speaks low and slow; weapons watch. Bulletin of the Atomic Scientists, 45(10):51, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNA

[NA89b] Robert S. Norris and William M. Arkin. Nuclear notebook: All in the family; missile mishaps; Navy freebies; INF watch. Bulletin of the Atomic Scientists, 45(7):47, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNB

[NA89c] Robert S. Norris and William M. Arkin. Nuclear notebook: Back to basics; where have all the soldiers gone?; all wearing sneakers?; the Iceworm cometh (and goeth); weapons watch. Bulletin of the Atomic Scientists, 45(6):55, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNC

[NA89d] Robert S. Norris and William M. Arkin. Nuclear notebook: Correction: [Chinese nuclear tests]. Bulletin of the Atomic Scientists, 45(9):52, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NA89g].

Norris:1989:NNE

[NA89e] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated Soviet nuclear stockpile, July 1989. Bulletin of the Atomic Scientists, 45(6):56, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNG

[NA89f] Robert S. Norris and William M. Arkin. Nuclear notebook: Good Nike SAM; arming with Tomahawks; burrowing in Nevada and Alaska; six-inch wonder; Soviet navy. Bulletin of the Atomic Scientists, 45(3):47, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNKb

[NA89g] Robert S. Norris and William M. Arkin. Nuclear notebook: Known Chinese nuclear tests, 1964–1988. Bulletin of the Atomic Scientists, 45(8):48, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [NA89d].

Norris:1989:NNKa

[NA89h] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945 to December 31, 1988. Bulletin of the Atomic Scientists, 45(3):48, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNL

[NA89i] Robert S. Norris and William M. Arkin. Nuclear notebook: Lulu’s grandchild due in 1993; costs of killing Spiders; two out of three is bad; B-2 secret: jobs, jobs, jobs. Bulletin of the Atomic Scientists, 45(8):47, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNMa

[NA89j] Robert S. Norris and William M. Arkin. Nuclear notebook: More bucks for the bang; 697 down, 1,995 to go [missiles destroyed]; weapons watch; Tridents on parade; Star Wars duty for Spartans. Bulletin of the Atomic Scientists, 45(2):51, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNNa

[NA89k] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear pursuits. Bulletin of the Atomic Scientists, 45(4):57, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SN89].

Norris:1989:NNNb

[NA89l] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear weapons at sea, 1989. Bulletin of the Atomic Scientists, 45(7):48, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNP

[NA89m] Robert S. Norris and William M. Arkin. Nuclear notebook: Pacific buildup; Trident trains; new war plan for 1990s; INF update. Bulletin of the Atomic Scientists, 45(5):48, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNS

[NA89n] Robert S. Norris and William M. Arkin. Nuclear notebook: Soviet strategic nuclear forces, end of 1988. Bulletin of the Atomic Scientists, 45(2):52, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNT

[NA89o] Robert S. Norris and William M. Arkin. Nuclear notebook: TAS-Mania; miscommunications Down Under; what’s new, pussycat?; weapons watch. Bulletin of the Atomic Scientists, 45(4):56, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNUb

[NA89p] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S.–Soviet nuclear weapons stockpile, 1945–1989: Numbers of weapons. Bulletin of the Atomic Scientists, 45(9):53, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNUa

[NA89q] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile (June 1989). Bulletin of the Atomic Scientists, 45(5):49, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNUc

[NA89r] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S., Soviet nuclear weapons stockpile, 1945–1989: Megatonnage. Bulletin of the Atomic Scientists, 45(10):52, December 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1989:NNW

[NA89s] Robert S. Norris and William M. Arkin. Nuclear notebook: What the Hades is going on?; missile ups and downs; perestroika watch; unfriendly skies; Nagasaki says “no thanks”. Bulletin of the Atomic Scientists, 45(9):52, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:BSS

[NA90a] Robert S. Norris and William M. Arkin. Beating swords into swords. Bulletin of the Atomic Scientists, 46(9):14–16, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://docs.nrdc.org/nuclear/nuc_

11019001a_022.pdf. See comment [Sut91].

Norris:1990:NNE

[NA90b] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated Soviet nuclear stockpile (July 1990). Bulletin of the Atomic Scientists, 46(6):49, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNH

[NA90c] Robert S. Norris and William M. Arkin. Nuclear notebook: Hot dogs; Trident limps on; missiles ship out; weapons watch. Bulletin of the Atomic Scientists, 46(10):56, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNJ

[NA90d] Robert S. Norris and William M. Arkin. Nuclear notebook: Just the fax, please; sparkling IDea; INF watch; Hermes flies; deadly submarine; Bears hibernating. Bulletin of the Atomic Scientists, 46(2):48, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNK

[NA90e] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945 to December 31, 1989. Bulletin of the Atomic Scientists, 46(3):57, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNLa

[NA90f] Robert S. Norris and William M. Arkin. Nuclear notebook: Lesser nuclear powers: Britain and China. Bulletin of the Atomic Scientists, 46(9):49, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNLb

[NA90g] Robert S. Norris and William M. Arkin. Nuclear notebook: Lesser nuclear powers: France. Bulletin of the Atomic Scientists, 46(10):57, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNNb

[NA90h] Robert S. Norris and William M. Arkin. Nuclear notebook: NATO nuclear weapons in Western Europe, 1990. Bulletin of the Atomic Scientists, 46(8):48–49, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNNa

[NA90i] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear weapons at sea, 1990. Bulletin of the Atomic Scientists, 46(7):49, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNNc

[NA90j] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear weapons in the Gulf; Soviet test numbers revealed; weapons watch. Bulletin of the Atomic Scientists, 46(9):48, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNO

[NA90k] Robert S. Norris and William M. Arkin. Nuclear notebook: Oakland may get its way after all; Turkey to lose some nuclear bombs; designer weapons for the 1990s; new bomb factory to open soon at test site. Bulletin of the Atomic Scientists, 46(3):56, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNR

[NA90l] Robert S. Norris and William M. Arkin. Nuclear notebook: Ready, set, START; SRAM recall; belt tightening; nuclear knowhow and an attitude; no deposit, no return. Bulletin of the Atomic Scientists, 46(7):48, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNSa

[NA90m] Robert S. Norris and William M. Arkin. Nuclear notebook: Soviet alerts; big, big cruisers; weapons watch; what is known about Israeli missiles. Bulletin of the Atomic Scientists, 46(1):48, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNSc

[NA90n] Robert S. Norris and William M. Arkin. Nuclear notebook: Soviet false alarms; inspecting Bears; special relationship; Europeans turn away from nuclear weapons. Bulletin of the Atomic Scientists, 46(5):47, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNSb

[NA90o] Robert S. Norris and William M. Arkin. Nuclear notebook: Soviet strategic nuclear forces, end of 1989. Bulletin of the Atomic Scientists, 46(2):49, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNTa

[NA90p] Robert S. Norris and William M. Arkin. Nuclear notebook: Then there were none; Soviet testing move; German Spiders spotted; axing Tomahawk? SRAM family. Bulletin of the Atomic Scientists, 46(4):56, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNTb

[NA90q] Robert S. Norris and William M. Arkin. Nuclear notebook: Trident 1.5? cruise news; Soviet bomber woes; close call; weapons watch. Bulletin of the Atomic Scientists, 46(6):48, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNUb

[NA90r] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile (June 1990). Bulletin of the Atomic Scientists, 46(5):48, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNUa

[NA90s] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. strategic nuclear forces, end of 1989. Bulletin of the Atomic Scientists, 46(1):49, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NNW

[NA90t] Robert S. Norris and William M. Arkin. Nuclear notebook: Wasting away. Bulletin of the Atomic Scientists, 46(4):57, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1990:NWG

[NA90u] Robert S. Norris and William M. Arkin. Nuclear weapons in the Gulf. Bulletin of the Atomic Scientists, 46(9):48, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:DRS

[NA91a] Robert S. Norris and William M. Arkin. Disarmament race, September–October 1991. Bulletin of the Atomic Scientists, 47 (10):49, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNB

[NA91b] Robert S. Norris and William M. Arkin. Nuclear notebook: Britain bomb papa; warhead orders flat; powering down; weapons watch. Bulletin of the Atomic Scientists, 47(4):48, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL https://en.wikipedia.org/wiki/William_
Penney,_Baron_Penney.

Norris:1991:NND

[NA91c] Robert S. Norris and William M. Arkin. Nuclear notebook: Disarmament race, September–October 1991. Bulletin of the Atomic Scientists, 47(10):49, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNE

[NA91d] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated Soviet nuclear stockpile (July 1991). Bulletin of the Atomic Scientists, 47(6):48, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNJ

[NA91e] Robert S. Norris and William M. Arkin. Nuclear notebook: Just curious; Tennessee fireworks; group therapy for Trident II. Bulletin of the Atomic Scientists, 47(6):47, July/August 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNK

[NA91f] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945 to December 31, 1990. Bulletin of the Atomic Scientists, 47(3):49, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNL

[NA91g] Robert S. Norris and William M. Arkin. Nuclear notebook: Latest Soviet test raises row; Eagles fly north; sticky rice; Energy Department headaches. Bulletin of the Atomic Scientists, 47(1): 47, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNMa

[NA91h] Robert S. Norris and William M. Arkin. Nuclear notebook: Maintenance is our profession; safety last. Bulletin of the Atomic Scientists, 47(2):48, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNMb

[NA91i] Robert S. Norris and William M. Arkin. Nuclear notebook: Museum pieces; uncoiling the spring; upset the fruit basket; weapons watch. Bulletin of the Atomic Scientists, 47(5):48, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNNb

[NA91j] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear pursuits. Bulletin of the Atomic Scientists, 47(4):49, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNNc

[NA91k] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear weapons at sea, 1990. Bulletin of the Atomic Scientists, 47(7):49, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNNa

[NA91l] Robert S. Norris and William M. Arkin. Nuclear notebook: “Nuke ’em”; nuclear ins and outs; farewell, Holy Loch. Bulletin of the Atomic Scientists, 47(3):48, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNP

[NA91m] Robert S. Norris and William M. Arkin. Nuclear notebook: Papa Teller says; if they exist, there are 100 of them; weapons watch; Silk Purse abandoned. Bulletin of the Atomic Scientists, 47(7):

48, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNS

[NA91n] Robert S. Norris and William M. Arkin. Nuclear notebook: Soviet strategic nuclear forces, end of 1990. Bulletin of the Atomic Scientists, 47(2):49, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNUb

[NA91o] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons safety and control features. Bulletin of the Atomic Scientists, 47(8):48–49, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNUa

[NA91p] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile (June 1991). Bulletin of the Atomic Scientists, 47(5):49, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1991:NNW

[NA91q] Robert S. Norris and William M. Arkin. Nuclear notebook: Where the weapons are. Bulletin of the Atomic Scientists, 47 (9):48–49, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNBa

[NA92a] Robert S. Norris and William M. Arkin. Nuclear notebook: B-52s, ALCMs “Desert Stormed” from Louisiana; gun control, sort of; an R2D2 for Pantex. Bulletin of the Atomic Scientists, 48 (3):48, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNBb

[NA92b] Robert S. Norris and William M. Arkin. Nuclear notebook: British–American TASMania. Bulletin of the Atomic Scientists, 48(9):47, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNC

[NA92c] Robert S. Norris and William M. Arkin. Nuclear notebook: C.I.S. (Soviet) strategic nuclear forces, end of 1991. Bulletin of the Atomic Scientists, 48(2):49, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNEb

[NA92d] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated CIS (Soviet) nuclear stockpile (July 1992). Bulletin of the Atomic Scientists, 48(6):49, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNEa

[NA92e] Robert S. Norris and William M. Arkin. Nuclear notebook: Every cloud has a silver lining; in the vanguard; beyond the fail-safe point; supersafe containers; all quiet on the French front. Bulletin of the Atomic Scientists, 48(6):48, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNF

[NA92f] Robert S. Norris and William M. Arkin. Nuclear notebook: Filling in the blanks: Operation Greenhouse; Operation Redwing; Operation Hardtack I. Bulletin of the Atomic Scientists, 48(9): 48, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNK

[NA92g] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945 to December 31, 1991. Bulletin of the Atomic Scientists, 48(3):49, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNPb

[NA92h] Robert S. Norris and William M. Arkin. Nuclear notebook: Pantex lays nukes to rest. Bulletin of the Atomic Scientists, 48(8): 48–49, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNPa

[NA92i] Robert S. Norris and William M. Arkin. Nuclear notebook: Proposed U.S. and C.I.S. strategic forces. Bulletin of the Atomic Scientists, 48(4):48–49, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano92-31].

Norris:1992:NNS

[NA92j] Robert S. Norris and William M. Arkin. Nuclear notebook: START extra; disarmament watch; who’s in charge? Bulletin of the Atomic Scientists, 48(1):48, January/February 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNM

[NA92k] Robert S. Norris and William M. Arkin. Nuclear notebook: The Midas touch; ferry warheads; last Blackjack; Chinese bomb pioneers; strike Eagles. Bulletin of the Atomic Scientists, 48(10): 57, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNU

[NA92l] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. strategic nuclear forces, end of 1991. Bulletin of the Atomic Scientists, 48(1):49, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:NNW

[NA92m] Robert S. Norris and William M. Arkin. Nuclear notebook: Where the bombs are. Bulletin of the Atomic Scientists, 48(7):49– 50, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:PLN

[NA92n] Robert S. Norris and William M. Arkin. Pantex lays nukes to rest. Bulletin of the Atomic Scientists, 48(8):48–49, October 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:TMM

[NA92o] Robert S. Norris and William M. Arkin. Tinynukes for mini minds. Bulletin of the Atomic Scientists, 48(3):24–25, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1992:UNW

[NA92p] Robert S. Norris and William M. Arkin. U.S. nuclear weapons stockpile (June 1992). Bulletin of the Atomic Scientists, 48(5): 48–49, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNB

[NA93a] Robert S. Norris and William M. Arkin. Nuclear notebook: Britain nuclear forces 1993. Bulletin of the Atomic Scientists, 49(7):57, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNC

[NA93b] Robert S. Norris and William M. Arkin. Nuclear notebook: Chinese nuclear forces 1993. Bulletin of the Atomic Scientists, 49(9): 57, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNEb

[NA93c] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated nuclear stockpiles 1945–1993. Bulletin of the Atomic Scientists, 49(10):57, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNEa

[NA93d] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated Russian (C.I.S.) nuclear stockpile (July 1993). Bulletin of the Atomic Scientists, 49(6):57, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNFa

[NA93e] Robert S. Norris and William M. Arkin. Nuclear notebook: Flying free — nuclear-free, that is. Bulletin of the Atomic Scientists, 49(8):55, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNSa

[NA93f] Robert S. Norris and William M. Arkin. Nuclear notebook: For sale-ski; Russia’s Pantexes; disarming, disarmingly; uranium glut. Bulletin of the Atomic Scientists, 49(1):56, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNFb

[NA93g] Robert S. Norris and William M. Arkin. Nuclear notebook: French nuclear forces 1993. Bulletin of the Atomic Scientists, 49 (8):56, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNK

[NA93h] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945 to December 31, 1992. Bulletin of the Atomic Scientists, 49(3):49, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ide93].

Norris:1993:NNNc

[NA93i] Robert S. Norris and William M. Arkin. Nuclear notebook: Naval base goes into dry dock; more warheads check out. Bulletin of the Atomic Scientists, 49(7):56, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNNb

[NA93j] Robert S. Norris and William M. Arkin. Nuclear notebook: Now is that clear?; through (with) the Looking Glass; NPT — onward and upward; Defensor Vindex redux; weapons watch. Bulletin of the Atomic Scientists, 49(6):56, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNNa

[NA93k] Robert S. Norris and William M. Arkin. Nuclear notebook: Nuclear pursuits. Bulletin of the Atomic Scientists, 49(4):48–49, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNP

[NA93l] Robert S. Norris and William M. Arkin. Nuclear notebook: Powering down quietly; have I got a deal for you. Bulletin of the Atomic Scientists, 49(5):56, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNRc

[NA93m] Robert S. Norris and William M. Arkin. Nuclear notebook: Royal commandos in the spotlight. Bulletin of the Atomic Scientists, 49 (8):55, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNRa

[NA93n] Robert S. Norris and William M. Arkin. Nuclear notebook: Russian (C.I.S.) strategic nuclear forces, end of 1992. Bulletin of the Atomic Scientists, 49(2):49, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNRb

[NA93o] Robert S. Norris and William M. Arkin. Nuclear notebook: Russian/Soviet weapons secrets revealed. Bulletin of the Atomic Scientists, 49(3):48, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-13610870.html.

Norris:1993:NNSb

[NA93p] Robert S. Norris and William M. Arkin. Nuclear notebook: Still going. Bulletin of the Atomic Scientists, 49(9):56, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNUc

[NA93q] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile (June 1993). Bulletin of the Atomic Scientists, 49(5):57, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNUa

[NA93r] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. strategic nuclear forces, end of 1992. Bulletin of the Atomic Scientists, 49(1):57, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNUb

[NA93s] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. weapons secrets revealed; Pantex takes it apart. Bulletin of the Atomic Scientists, 49(2):48, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:SG

[NA93t] Robert S. Norris and William M. Arkin. Still going. Bulletin of the Atomic Scientists, 49(9):56, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNEb

[NA94a] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated U.S. and Soviet/Russian nuclear stockpiles, 1945–94. Bulletin of the Atomic Scientists, 50(6):58–59, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNK

[NA94b] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945–1993. Bulletin of the Atomic Scientists, 50(3):62–63, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNO

[NA94c] Robert S. Norris and William M. Arkin. Nuclear notebook: Operational U.S. nuclear weapons stockpile, July 1994. Bulletin of the Atomic Scientists, 50(4):63, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNR

[NA94d] Robert S. Norris and William M. Arkin. Nuclear notebook: Russian (C.I.S.) strategic nuclear forces: End of 1993. Bulletin of the Atomic Scientists, 50(2):63, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNNb

[NA94e] Robert S. Norris and William M. Arkin. Nuclear notebook: The next generation; GWEN will I see you again?; highly enriched. Bulletin of the Atomic Scientists, 50(2):62, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNNa

[NA94f] Robert S. Norris and William M. Arkin. Nuclear notebook: To nuke or not to nuke; new base in the Mediterranean? Bulletin of the Atomic Scientists, 50(1):64, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNUb

[NA94g] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile, July 1994. Bulletin of the Atomic Scientists, 50(4):61–63, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNUa

[NA94h] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. strategic nuclear forces, end of 1993. Bulletin of the Atomic Scientists, 50(1):65, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1995:NNN

[NA95a] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. nuclear weapon locations, 1995. Bulletin of the Atomic Scientists, 51(6):74–75, November/December 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1995:NNE

[NA95b] Robert S. Norris and William M. Arkin. Nuclear notebook: Estimated Russian (C.I.S.) stockpile, September 1995. Bulletin of the Atomic Scientists, 51(5):62–63, 61, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1995:NNK

[NA95c] Robert S. Norris and William M. Arkin. Nuclear notebook: Known nuclear tests worldwide, 1945–1994. Bulletin of the Atomic Scientists, 51(3):70–71, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1995:NNR

[NA95d] Robert S. Norris and William M. Arkin. Nuclear notebook: Russian (C.I.S.) strategic nuclear forces. Bulletin of the Atomic Scientists, 51(2):78–79, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1995:NNU

[NA95e] Robert S. Norris and William M. Arkin. Nuclear notebook: U.S. nuclear weapons stockpile, July 1995. Bulletin of the Atomic Scientists, 51(4):77–79, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1995:USN

[NA95f] Robert S. Norris and William M. Arkin. U.S. strategic nuclear forces, end of 1994. Bulletin of the Atomic Scientists, 51(1):69– 71, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1996:NNNf

[NA96a] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: British, French, and Chinese nuclear forces. Bulletin of the Atomic Scientists, 52(6):64–67, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1996:NNNe

[NA96b] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Estimated Russian stockpile, September 1996. Bulletin of the Atomic Scientists, 52(5):62–63, September/October 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1996:NNNc

[NA96c] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Known nuclear tests worldwide, 1945–1995. Bulletin of the Atomic Scientists, 52(3):61–63, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1996:NNNb

[NA96d] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Russian (C.I.S.) strategic nuclear forces end of 1995. Bulletin of the Atomic Scientists, 52(2):62–63, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1996:NNNd

[NA96e] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. nuclear weapons stockpile, July 1996. Bulletin of the Atomic Scientists, 52(4):61–63, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1996:NNNa

[NA96f] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. strategic nuclear forces, end of 1995. Bulletin of the Atomic Scientists, 52(1):62–63, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1997:NNNc

[NA97a] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Estimated Russian stockpile, end of 1996. Bulletin of the Atomic Scientists, 53(3):62–64, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1997:NNNf

[NA97b] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Global nuclear stockpiles, 1945–1997. Bulletin of the Atomic Scientists, 53(6):67, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1997:NNNb

[NA97c] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: The Internet and the bomb. Bulletin of the Atomic Scientists, 53 (2):62–63, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1997:NNNd

[NA97d] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. nuclear stockpile, July 1997. Bulletin of the Atomic Scientists, 53(4):62–63, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1997:NNNa

[NA97e] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. strategic nuclear forces, end of 1996. Bulletin of the Atomic Scientists, 53(1):70–71, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1997:NNNe

[NA97f] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Where the bombs are, 1997. Bulletin of the Atomic Scientists, 53(5):62–63, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:ATI

[NA98a] Robert S. Norris and William M. Arkin. After the test: India and Pakistan update. Bulletin of the Atomic Scientists, 54(5):69, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:NNNe

[NA98b] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: After the tests: India and Pakistan update. Bulletin of the Atomic Scientists, 54(5):69–71, September/October 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:NNNf

[NA98c] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Known nuclear tests worldwide, 1945–98. Bulletin of the Atomic Scientists, 54(6):65–67, November/December 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:NNNb

[NA98d] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Russian strategic nuclear forces, end of 1997. Bulletin of the Atomic Scientists, 54(2):70–71, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:NNNc

[NA98e] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: Soviet nuclear testing, August 29, 1949–October 24, 1990. Bulletin of the Atomic Scientists, 54(3):69–71, May/June 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:NNNd

[NA98f] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. nuclear stockpile, July 1998. Bulletin of the Atomic Scientists, 54(4):69–71, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1998:NNNa

[NA98g] Robert S. Norris and William M. Arkin. NRDC nuclear notebook: U.S. strategic nuclear forces, end of 1997. Bulletin of the Atomic Scientists, 54(1):70–72, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1999:CNF

[NA99a] Robert S. Norris and William M. Arkin. Chinese nuclear forces, 1999. Bulletin of the Atomic Scientists, 55(3):79–80, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/3/79.

Norris:1999:FBN

[NA99b] Robert S. Norris and William M. Arkin. French and British nuclear forces, 1999. Bulletin of the Atomic Scientists, 55(4): 77–79, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/4/77.

Norris:1999:RSN

[NA99c] Robert S. Norris and William M. Arkin. Russian strategic nuclear forces, end of 1998. Bulletin of the Atomic Scientists, 55(2):62–63, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
2/62.

Norris:1999:USN

[NA99d] Robert S. Norris and William M. Arkin. U.S. strategic nuclear forces, end of 1998. Bulletin of the Atomic Scientists, 55(1):78– 80, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/55/1/78.

Norris:1999:WPI

[NA99e] Robert S. Norris and William M. Arkin. World plutonium inventories. Bulletin of the Atomic Scientists, 55(5):71, September/ October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/55/
5/71.

Norris:2000:CNF

[NA00a] Robert S. Norris and William M. Arkin. Chinese nuclear forces, 2000. Bulletin of the Atomic Scientists, 56(6):78–79, November/ December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
6/78.

Norris:2000:FBN

[NA00b] Robert S. Norris and William M. Arkin. French and British nuclear forces, 2000. Bulletin of the Atomic Scientists, 56(5):69–71, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/5/69.

Norris:2000:GNS

[NA00c] Robert S. Norris and William M. Arkin. Global nuclear stockpiles, 1945–2000. Bulletin of the Atomic Scientists, 56(2):79, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/2/79.

Norris:2000:HMD

[NA00d] Robert S. Norris and William M. Arkin. How much did Japan know? Bulletin of the Atomic Scientists, 56(1):11, January/ February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:2000:RNF

[NA00e] Robert S. Norris and William M. Arkin. Russian nuclear forces, 2000. Bulletin of the Atomic Scientists, 56(4):70–71, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/4/70.

Norris:2000:UNF

[NA00f] Robert S. Norris and William M. Arkin. U.S. nuclear forces, 2000. Bulletin of the Atomic Scientists, 56(3):69–71, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/3/69.

Norris:2001:NSW

[NA01] Robert S. Norris and William M. Arkin. National security on the Web. Bulletin of the Atomic Scientists, 57(1):74–75, January/ February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
1/74.

Norris:1999:ABD

[NAB99a] Robert S. Norris, William M. Arkin, and William Burr. Appendix B: Deployments by country, 1951–1977. Bulletin of the Atomic Scientists, 55(6):66–67, November/December 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/6/66.

Norris:1999:WTW

[NAB99b] Robert S. Norris, William M. Arkin, and William Burr. Where they were. Bulletin of the Atomic Scientists, 55 (6):26–35, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:/

/bos.sagepub.com/content/55/6/26;   http://docs.nrdc.
org/nuclear/nuc_11019901a_001.pdf.

Norris:2000:WTW

[NAB00] Robert S. Norris, William M. Arkin, and William Burr. Where they were: How much did Japan know? Bulletin of the Atomic Scientists, 56(1):11–13, January/February 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/1/11;   http://
docs.nrdc.org/nuclear/nuc_00010101a_002.pdf.

Nace:1965:ICS

[Nac65] Raymond L. Nace. International cooperation in science: Water: A common problem. Bulletin of the Atomic Scientists, 21(1):32– 34, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nace:1969:ATL

[Nac69] Raymond L. Nace. Arrogance toward the landscape: A problem of water planning. Bulletin of the Atomic Scientists, 25(10):11– 14, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nace:1972:MWL

[Nac72] Raymond L. Nace. Man and water: A lesson in history. Bulletin of the Atomic Scientists, 28(3):34–38, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1987:UUS

[NAC87] Robert S. (Robert Stan) Norris, William M. Arkin, and Thomas B. Cochran. US–USSR strategic offensive nuclear forces, 1946–1987. Nuclear weapons databook working papers. Natural Resources Defense Council, Washington, DC, USA, 1987. ???? pp. LCCN U264.3 .N67 1987. NWD 87-1 (Rev. 1). December 1987.

Nichols:1947:SPD

[NAD^{+}47] General Kenneth D. Nichols, Captain Alvara Alberto, Sir Charles Darwin, Ralph L. Harry, General Andrew G. L. McNaughton, and François de Rose. Stockpiling, production and distribution of nuclear fuels. Bulletin of the Atomic Scientists, 3(9):260–261, 266, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nader:1966:TED

[Nad66] Claire Nader. The technical expert in a democracy. Bulletin of the Atomic Scientists, 22(5):28–30, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nader:1972:SHI

[Nad72] Ralph Nader. The scientist and his indentured professional societies. Bulletin of the Atomic Scientists, 28(2):43–44, 46, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nader:1979:BRB

[Nad79] Laura Nader. Book reviews: Scientific Elite: Nobel Laureates in the United States, by Harriet Zuckerman. Bulletin of the Atomic Scientists, 35(1):40–41, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nadis:1980:TR

[Nad80] Steven J. Nadis. Time for a reassessment. Bulletin of the Atomic Scientists, 36(2):37–44, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nadal:1998:BRG

[Nad98a] Alejandro Nadal. Book review: Globalism v. democracy: Global Squeeze: The Coming Crisis for First World Nations, by Richard C. Longworth. Bulletin of the Atomic Scientists, 54(5):64–66, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nadal:1998:PCC

[Nad98b] Alejandro Nadal. Political cleansing in Chiapas. Bulletin of the Atomic Scientists, 54(3):20–22, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nadal:1998:TC

[Nad98c] Alejandro Nadal. Terror in Chiapas. Bulletin of the Atomic Scientists, 54(2):18–25, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Han98b].

Nadal:1998:TLP

[Nad98d] Alejandro Nadal. Trashing the “Law for Peace”. Bulletin of the Atomic Scientists, 54(2):25, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nadal:2004:MWT

[Nad04] Alejandro Nadal. Mexico: Worse than you know. Bulletin of the Atomic Scientists, 60(2):27–29, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/27.

Nagle:1971:GE

[Nag71] James J. Nagle. Genetic engineering. Bulletin of the Atomic Scientists, 27(10):43–45, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nagler:1981:CPP

[Nag81] Michael N. Nagler. Commentary: Peace as a paradigm shift. Bulletin of the Atomic Scientists, 37(10):49–52, December 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nagler:1984:RP

[Nag84] Michael Nagler. Redefining peace. Bulletin of the Atomic Scientists, 40(9):36–38, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hun85c].

Nagappa:2015:GTF

[Nag15] Rajaram Nagappa. Going too fast: Time to ban hypersonic missile tests? : an Indian response. Bulletin of the Atomic Scientists, 71(5):9–12, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/9.full.pdf+html.

Naidoo:1981:CVC

[Nai81] L. Naidoo. Commentary: Vital concerns in South Africa. Bulletin of the Atomic Scientists, 37(10):58, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:2001:BNF

[NAKH01a] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. British nuclear forces, 2001. Bulletin of the Atomic Scientists, 57(6):78–79, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/6/78.

Norris:2001:CNF

[NAKH01b] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. Chinese nuclear forces, 2001. Bulletin of the Atomic Scientists, 57(5):71–72, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/71.

Norris:2001:FNF

[NAKH01c] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. French nuclear forces, 2001. Bulletin of the Atomic Scientists, 57(4):70–71, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/4/70.

Norris:2001:RNF

[NAKH01d] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. Russian nuclear forces, 2001. Bulletin of the Atomic Scientists, 57(3):78–79, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/78.

Norris:2001:UNF

[NAKH01e] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. U.S. nuclear forces, 2001. Bulletin of the Atomic Scientists, 57(2):77–79, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/2/77.

Norris:2002:INFa

[NAKH02a] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. India’s nuclear forces, 2002. Bulletin of the Atomic Scientists, 58(2):70–72, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/70.

Norris:2002:INFb

[NAKH02b] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. Israeli nuclear forces, 2002. Bulletin of the Atomic Scientists, 58(5):73–75, September/October 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/5/73.

Norris:2002:PNF

[NAKH02c] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. Pakistan’s nuclear forces, 2001. Bulletin of the Atomic Scientists, 58(1):70–71, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/70.

Norris:2002:RNF

[NAKH02d] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. Russian nuclear forces, 2002. Bulletin of the Atomic Scientists, 58(4):71–73, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/4/71.

Norris:2002:UNF

[NAKH02e] Robert S. Norris, William M. Arkin, Hans M. Kristensen, and Joshua Handler. U.S. nuclear forces, 2002. Bulletin of the Atomic Scientists, 58(3):70–75, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/3/70.

Nanda:1971:WIG

[Nan71] Krish Nanda. Will India go nuclear? Bulletin of the Atomic Scientists, 27(10):39–41, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nanda:1973:BBC

[Nan73] Krish Nanda. Books: Protest in City Politics: Rent Strikes, Housing and the Power of the Poor, by Michael Lipsky. Bulletin of the Atomic Scientists, 29(2):46–48, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nansen:1978:P

[Nan78] Fridtjof Nansen. Perspective: 1922. Bulletin of the Atomic Scientists, 34(7):1, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the author’s Nobel Prize Lecture, 29 December 1922, Oslo, Norway.

Narayan:1967:CBB

[Nar67] Shri Jayaprakash Narayan. Comments: Building the bomb: An Indian view. Bulletin of the Atomic Scientists, 23(2):25–26, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nash:1962:BBR

[Nas62] June Nash. Books: The Rich Nations and the Poor Nations, by Barbara Ward. Bulletin of the Atomic Scientists, 18(9):26, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nash:1980:BH

[Nas80] Henry T. Nash. The bureaucratization of homicide. Bulletin of the Atomic Scientists, 36(4):22–27, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nash:1982:BRB

[Nas82] Henry T. Nash. Book review: The Defense Industry, by Jacques S. Gansler. Bulletin of the Atomic Scientists, 38(7):39–40, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nash:1983:TAT

[Nas83] Henry T. Nash. Thinking about thinking about the unthinkable. Bulletin of the Atomic Scientists, 39(8):39–42, October 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SWC^{+}84].

Nasir:2003:AMI

[Nas03a] Sohail Abdul Nasir. Afghanistan: The more it changes. … Bulletin of the Atomic Scientists, 59(2):40–48, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/2/40.

Nasir:2003:AQT

[Nas03b] Sohail Abdul Nasir. Al Qaeda, two years on. Bulletin of the Atomic Scientists, 59(5):??, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/5/32.

Nasir:2004:APP

[Nas04a] Sohail Abdul Nasir. Afghanistan: The poppy problem. Bulletin of the Atomic Scientists, 60(5):15–16, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/5/15.

Nasir:2004:BDM

[Nas04b] Sohail Abdul Nasir. Bungling democracy in the Muslim world. Bulletin of the Atomic Scientists, 60(1):28–30, January 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/1/28.

Nash:2006:CS

[Nas06] Philip Nash. Cloak and swagger. Bulletin of the Atomic Scientists, 62(5):62–63, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/62.

Nash:2018:SRW

[Nas18] Stephen Nash. As seas rise, world awaits “the plan”. Bulletin of the Atomic Scientists, 74(3):160–161, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1461971.

NPA:1955:APN

[Nat55] National Planning Association. American policy in the new phase of the Cold War. Bulletin of the Atomic Scientists, 11(3):97–99, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

NAS:1956:LUR

[Nat56] National Academy of Sciences. Loyalty and unclassified research. Bulletin of the Atomic Scientists, 12(6):227–228, June 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathan:1963:EPP

[Nat63] Otto Nathan. The economics of permanent peace. Bulletin of the Atomic Scientists, 19(6):21–24, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathan:1967:CSC

[Nat67] Otto Nathan. Comments: The Sobell case. Bulletin of the Atomic Scientists, 23(4):36, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathan:1968:NF

[Nat68a] Otto Nathan. Nathan on Feld. Bulletin of the Atomic Scientists, 24(7):39, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel68d].

NSC:1968:NSC

[Nat68b] National Security Council. National Security Council on the draft. Bulletin of the Atomic Scientists, 24(4):39, April 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathan:1974:LRE

[Nat74] Otto Nathan. Letter: A renewed effort. Bulletin of the Atomic Scientists, 30(10):2–3, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Can74].

NSG:1977:PME

[Nat77] Natural Science Group. The physical and medical effects of the Hiroshima and Nagasaki bombs. Bulletin of the Atomic Scientists, 33(10):54–56, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste78].

Nathan:1981:DPO

[Nat81] Ove Nathan. Denmark: power options still open. Bulletin of the Atomic Scientists, 37(3):29–33, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathanson:1984:POA

[Nat84] Melvyn B. Nathanson. Perspective: Orwell and the atom bomb. Bulletin of the Atomic Scientists, 40(4):4, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathanson:1985:SAF

[Nat85a] Melvyn B. Nathanson. Secrecy: Academic freedom versus nonproliferation: the Libyan case. Bulletin of the Atomic Scientists, 41(3):29–31, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathanson:1985:SRO

[Nat85b] Melvyn B. Nathanson. Soviet reactors to open for international inspection. Bulletin of the Atomic Scientists, 41(6):32–33, June/ July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

NRC:1985:NWR

[Nat85c] National Research Council. Nuclear winter report excerpts. Bulletin of the Atomic Scientists, 41(3):39–40, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathan:1986:PCS

[Nat86a] Ove Nathan. Perspective: A call for Soviet operations. Bulletin of the Atomic Scientists, 42(8):8, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathanson:1986:LSN

[Nat86b] Melvyn B. Nathanson. Letter: Success at the NPT review conference. Bulletin of the Atomic Scientists, 42(2):48, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AC85] and comment [Loe86].

Nathanson:1987:BRB

[Nat87] Stephen Nathanson. Book review: Nuclear Ethics, by Joseph S. Nye, Jr. Bulletin of the Atomic Scientists, 43(1):53–54, January/ February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nathan:1988:ASD

[Nat88] Ove Nathan. After the summit: Danes look to détente for greater security. Bulletin of the Atomic Scientists, 44(2):32–34, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nayan:2012:HCC

[Nay12] Rajiv Nayan. How can countries ensure that the nuclear security summit does not lose momentum and become just another gathering?: Position: Consensus, not enforcement. Bulletin of the Atomic Scientists, 68(2):84–86, March/April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/68/2/84.

Noel-Baker:1977:POI

[NB77] Philip Noel-Baker. Peace or oblivion? an inescapable choice in the atomic age. Bulletin of the Atomic Scientists, 33(7):18–20, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Cultures 3(4) 1976. See comment [Hue77].

Nathan:1978:SSQ

[NB78] Leonard Nathan and William J. Brandt. Scientists speak with a ‘quarked’ tongue. Bulletin of the Atomic Scientists, 34(4):22–23, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Noel-Baker:1979:PNP

[NB79] Philip Noel-Baker. Perspective: Nobel Peace Laureates: 1959. Bulletin of the Atomic Scientists, 35(5):1, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 10 December 1959, Oslo, Norway.

Norris:1994:NWD

[NBF94] Robert S. (Robert Stan) Norris, Andrew S. Burrows, and Richard W. Fieldhouse. Nuclear weapons databook. Volume 5. British, French, and Chinese nuclear weapons, volume 5 of Nuclear weapons databook. Westview Press, Boulder, CO, USA, 1994. ISBN 0-8133-1611-1 , 0-8133-1612-X (paperback). xvii + 437 pp. LCCN U264 .N827 1984 vol. 5.

Neel:1985:DBE

[NBM85] James V. Neel, Gilbert W. Beebe, and Robert W. Miller. Delayed biomedical effects of the bombs. Bulletin of the Atomic Scientists, 41(7):72–75, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:2006:CNF

[NC06] Robert S. Norris and Hans M. Christensen. Chinese nuclear forces, 2006. Bulletin of the Atomic Scientists, 62(3):60–63, May/ June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/62/3/60.

Norris:1985:HNS

[NCA85] Robert S. Norris, Thomas B. Cochran, and William M. Arkin. History of the nuclear stockpile. Bulletin of the Atomic Scientists, 41(7):106–109, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://docs.nrdc.org/

nuclear/nuc_85080001a_55.pdf.

Norris:1988:KUN

[NCA88a] Robert S. Norris, Thomas B. Cochran, and William M. Arkin. Known U.S. nuclear tests July 1945 to 31 December 1987, volume NWD 86-2 of Nuclear weapons databook working papers. Natural Resources Defense Council, Washington, DC, USA, 1988. 60 pp. LCCN U264 .N67 1988.

Norris:1988:WRB

[NCA88b] Robert S. Norris, Thomas B. Cochran, and William M. Arkin. Washington report: Behind the disclosure of US secret tests. Bulletin of the Atomic Scientists, 44(2):3–4, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://docs.nrdc.org/nuclear/nuc_88030001a_81.pdf.

Norwood:1976:LPT

[NE76] W. D. Norwood and John T. Edsall. Letter: Plutonium toxicity data misinterpreted. Bulletin of the Atomic Scientists, 32(10): 6, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds76].

Neeman:1981:CIF

[Ne’81] Yuval Ne’eman. Commentary from Israel: The Franco–Iraqi project. Bulletin of the Atomic Scientists, 37(7):8–10, August/ September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fei81c, WK82].

Neal:1960:CK

[Nea60] Fred Warner Neal. Coexistence and the Kremlin. Bulletin of the Atomic Scientists, 16(7):283–288, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neal:1961:LEP

[Nea61] Fred Warner Neal. Letter to the Editor: Plan for Berlin. Bulletin of the Atomic Scientists, 17(3):121, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neal:1963:UCP

[Nea63] Fred Warner Neal. U.S. China policy and disarmament. Bulletin of the Atomic Scientists, 19(9):5–8, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neal:1978:AWDa

[Nea78a] Fred Warner Neal. Around the world in 60 days: a political travel diary. Bulletin of the Atomic Scientists, 34(8):27–31, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neal:1978:AWDb

[Nea78b] Fred Warner Neal. Around the world in 60 days: a political travel diary: Part 2. Bulletin of the Atomic Scientists, 34(9):29– 33, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neal:1980:CAC

[Nea80] Fred Warner Neal. Commentary: Afghanistan: a created crisis. Bulletin of the Atomic Scientists, 36(8):10–11, October 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neack:2002:GTO

[Nea02] Laura Neack. A government of their own. Bulletin of the Atomic Scientists, 58(6):72–80, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/6/72.

Neack:2004:PBP

[Nea04] Laura Neack. Peacekeeping, bloody peacekeeping. Bulletin of the Atomic Scientists, 60(4):40–47, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/40.

Needham:1979:GLF

[Nee79] Joseph Needham. Great leap forward means more contact with the West. Bulletin of the Atomic Scientists, 35(1):54–56, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Nature.

Nef:1959:ASL

[Nef59] John Nef. Art, science, and life. Bulletin of the Atomic Scientists, 15(2):86–88, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neff:1978:BBE

[Nef78] Thomas L. Neff. Books: Soft Energy Paths: Toward a Durable Peace, by Amory B. Lovins. Bulletin of the Atomic Scientists, 34 (6):58–61, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neier:1982:BRB

[Nei82a] Aryeh Neier. Book reviews: Calling a Truce to Terror, by Ernest Evans; Studies in Nuclear Terrorism, edited by Augustus R. Norton and Martin Greenberg; Taking Lives, by Irving Louis Horowitz; Ten Years of Terrorism; Terrorism: Theory and Practice, edited by Yonah Alexander, David Carlton, and Paul Wilkinson. Bulletin of the Atomic Scientists, 38(3):31–33, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neild:1982:HMY

[Nei82b] Robert Neild. How to make up your mind about the bomb. Bulletin of the Atomic Scientists, 38(1):13–18, January 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neild:1988:DOF

[Nei88] Robert Neild. Defining “offensive”: a failure and a success. Bulletin of the Atomic Scientists, 44(7):18, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1957:NTE

[Nel57] Thomas Nelson. Nuclear technology: An economic catalyst: I. The Cold War and foreign investment. Bulletin of the Atomic Scientists, 13(5):157–161, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1964:ARD

[Nel64] Richard R. Nelson. Adjusting R and D. Bulletin of the Atomic Scientists, 20(4):15–19, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelkin:1974:BBE

[Nel74a] Dorothy Nelkin. Books: Energy, Ecology, Economy: A Framework for Environmental Policy, by Gerald Garvey. Bulletin of the Atomic Scientists, 30(2):47–48, February 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelkin:1974:REN

[Nel74b] Dorothy Nelkin. The role of experts in a nuclear siting controversy: New York case pitted academe against industry. Bulletin of the Atomic Scientists, 30(9):29–36, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1975:BAS

[Nel75] Bryce Nelson. Bulletin of the Atomic Scientists: Thirty years of clockwatching. Science, 190(4219):1070–1073, December 12, 1975. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1741225.

Nelson:1977:CPN

[Nel77] Bryce Nelson. Corn patch Nobel laureate. Bulletin of the Atomic Scientists, 33(8):48–50, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelkin:1981:ANC

[Nel81] Dorothy Nelkin. Anti-nuclear connections: power and weapons. Bulletin of the Atomic Scientists, 37(4):36–40, April 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MD81].

Nelson:1984:BNB

[Nel84] Dennis R. Nelson. Book note: Brittle Power: Energy Strategy for National Security, by Amory B. Lovins and L. Hunter Lovins. Bulletin of the Atomic Scientists, 40(3):48, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1990:AQE

[Nel90] Daniel N. Nelson. Not all quiet on the Eastern Front. Bulletin of the Atomic Scientists, 46(9):35–38, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1992:PNM

[Nel92a] Daniel N. Nelson. Perspective: NATO — means, but no ends. Bulletin of the Atomic Scientists, 48(1):10–11, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1992:SDJ

[Nel92b] Daniel N. Nelson. Snatching defeat from the jaws of victory. Bulletin of the Atomic Scientists, 48(8):24–28, October 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1992:RNT

[Nel92c] Eric Nelson. Report: Nuclear testing: Britain’s aboriginal sin. Bulletin of the Atomic Scientists, 48(6):8–9, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1993:AEM

[Nel93a] Daniel N. Nelson. Ancient enmities, modern guns. Bulletin of the Atomic Scientists, 49(10):21–27, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1993:RHS

[Nel93b] Eric Nelson. Report: Hanford: Startup is cleanup, says Energy [Department]. Bulletin of the Atomic Scientists, 49(10):6–7, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1994:NUO

[Nel94] Daniel N. Nelson. NATO: Use only in moderation. Bulletin of the Atomic Scientists, 50(6):32–35, 60, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelson:1999:DC

[Nel99] Daniel N. Nelson. Damage control. Bulletin of the Atomic Scientists, 55(1):55–57, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/55.

Nelson:2000:DA

[Nel00] Daniel N. Nelson. Dangerous assumptions. Bulletin of the Atomic Scientists, 56(4):24–27, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/4/24.

Nelson:2009:RWU

[Nel09] Robert Nelson. 3 reasons why the U.S. Senate should ratify the Test Ban Treaty. Bulletin of the Atomic Scientists, 65(2): 52–58, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/65/2/52.

Nelson:2010:RNR

[Nel10] Paul Nelson. Reassessing the nuclear renaissance. Bulletin of the Atomic Scientists, 66(4):11–22, July/August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/66/4/11.

Nelson:2012:TNT

[Nel12] Andrew T. Nelson. Thorium: Not a near-term commercial nuclear fuel. Bulletin of the Atomic Scientists, 68(5):33–44, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/5/33.

Neri:1969:CBM

[Ner69] Umberto Neri. Communications: Black magic in science? Bulletin of the Atomic Scientists, 25(5):49, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro68, Bro69].

Nero:1976:LUS

[Ner76] Anthony V. Nero, Jr. Letter: Uranium ‘straw man’. Bulletin of the Atomic Scientists, 32(5):2, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75a, Fis76].

Nero:1980:BLW

[Ner80] Anthony V. Nero, Jr. Beyond the light water reactor. Bulletin of the Atomic Scientists, 36(2):31–37, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Netschert:1969:ATE

[Net69] Bruce C. Netschert. Antipollution technology: The electric car. Bulletin of the Atomic Scientists, 25(4):37, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Netschert:1970:EIE

[Net70] Bruce C. Netschert. The economic impact of electric vehicles: A scenario. Bulletin of the Atomic Scientists, 26(5):29–35, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jon71].

Netschert:1971:ECM

[Net71] Bruce C. Netschert. The energy company: A monopoly trend in the energy markets. Bulletin of the Atomic Scientists, 27(8): 13–17, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Netschert:1975:UCF

[Net75] Bruce C. Netschert. Use of capacity factor in the nuclear debate. Bulletin of the Atomic Scientists, 31(8):42–43, October 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com75g].

Neuman:1958:RMB

[Neu58a] W. F. Neuman. Radiation and man: Basic data: The somatic effects of fission production. Bulletin of the Atomic Scientists, 14(1):15–18, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neuman:1958:RMN

[Neu58b] W. F. Neuman. Radiation and man: Nuclear tests and radiation: Uncertainties in evaluating the effects of fall-out from weapons tests. Bulletin of the Atomic Scientists, 14(1):31–34, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neureiter:1965:UJC

[Neu65] Norman P. Neureiter. U.S.–Japan cooperative science program. Bulletin of the Atomic Scientists, 21(1):39, January 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neufeld:2007:WBU

[Neu07] Michael J. Neufeld. Wernher von Braun’s ultimate weapon. Bulletin of the Atomic Scientists, 63(4):??, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/4/50.

Newman:1954:TAA

[New54] James R. Newman. Toward atomic agreement. Bulletin of the Atomic Scientists, 10(4):121–122, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Newell:1958:SRC

[New58] Homer E. Newell, Jr. Space research — a challenge to mankind. Bulletin of the Atomic Scientists, 14(5):177–179, May 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Newell:1961:NS

[New61] Homer E. Newell. NASA and space. Bulletin of the Atomic Scientists, 17(5–6):222–229, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

NYCLACCR:1962:CC

[New62a] New York County Lawyers’ Association Committee on Civil Rights. Censorship on the campus. Bulletin of the Atomic Scientists, 18(9):46–48, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Newman:1962:NRC

[New62b] Stanley L. Newman. News & reports: Civil defense and the Congress: Quiet reversal. Bulletin of the Atomic Scientists, 18 (9):33–38, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Newhouse:1964:FMF

[New64] John Newhouse. Forum: The multilateral force: An appraisal. Bulletin of the Atomic Scientists, 20(7):13–18, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Newcombe:1967:CCP

[New67] Hanna Newcombe. Comments: The Canadian Peace Research Institute. Bulletin of the Atomic Scientists, 23(4):36–37, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rin66].

Newill:1973:PPC

[New73] Vaun A. Newill. Pollution’s price — the cost in human health. Bulletin of the Atomic Scientists, 29(8):47–49, October 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Neyman:1969:CCS

[Ney69] Jerzy Neyman. Current comments: Science and politics of rainmaking: A rejoinder. Bulletin of the Atomic Scientists, 25(3): 27, 32, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mac68a].

Neyman:1973:SBS

[Ney73] Jerzy Neyman. Sixth Berkeley Symposium on Mathematical Statistics and Probability: Epilogue of the health-pollution conference. Bulletin of the Atomic Scientists, 29(7):25–34, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Newby-Fraser:1983:CSA

[NF83] A. R. Newby-Fraser. Commentary: South Africa’s nuclear program. Bulletin of the Atomic Scientists, 39(4):47–48, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ade82].

Nasu:2013:NJR

[NF13] Hitoshi Nasu and Thomas Faunce. Nanotechnology in Japan: a route to energy security after Fukushima? Bulletin of the Atomic Scientists, 69(5):68–74, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/5/68.

Noggerath:2011:FMS

[NGG11] Johannis Nöggerath, Robert J. Geller, and Viacheslav K. Gusiakov. Fukushima: The myth of safety, the reality of geoscience. Bulletin of the Atomic Scientists, 67(5):37–46, September/ October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
5/37.

Nolan:2008:BD

[NH08a] Janne E. Nolan and James R. Holmes. The bureaucracy of deterrence. Bulletin of the Atomic Scientists, 64(1):40–43, March/ April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/64/1/40.

Nolan:2008:BDR

[NH08b] Janne E. Nolan and James R. Holmes. The bureaucracy of deterrence: To remake U.S. nuclear weapons policy, the next president will need to overcome entrenched interests. How arduous a task will this be? Ask Bill Clinton. Bulletin of the Atomic Scientists, 64(1):40–58, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nicholas:1954:BBW

[Nic54a] Herbert G. Nicholas. Books: Strategy for the West, by Sir John Slessor. Bulletin of the Atomic Scientists, 10(9):363–364, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nichols:1954:NPC

[Nic54b] Major General K. D. Nichols. Nichols presents charges [against J. Robert Oppenheimer]. Bulletin of the Atomic Scientists, 10(5): 174–176, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Opp54b].

Nichols:2023:NPC

[Nic23] K. D. Nichols. Nichols presents charges. Bulletin of the Atomic Scientists, 79(4):238–241, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2223083.

Niebuhr:1949:IWG

[Nie49] Reinhold Niebuhr. The illusion of world government. Bulletin of the Atomic Scientists, 5(10):289–292, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Niebuhr:1955:LSM

[Nie55] Reinhold Niebuhr. Limitations of the scientific method: An answer to Pierre Auger. Bulletin of the Atomic Scientists, 11(3):87, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Aug55] and comment [SD55].

Nieburg:1963:BBN

[Nie63] Harold L. Nieburg. Books: The New World, 1936–1946, Volume I of a History of the United States Atomic Energy Commission, by Richard G. Hewlett and Oscar E. Anderson, Jr. Bulletin of the Atomic Scientists, 19(1):29–31, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nierenberg:1965:NSP

[Nie65] William A. Nierenberg. The Nato Science Program. Bulletin of the Atomic Scientists, 21(5):45–48, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nieburg:1966:FRC

[Nie66a] H. L. Nieburg. Forum: R&D and the contract state: Throwing away the yardstick. Bulletin of the Atomic Scientists, 22(3):20– 24, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BN67].

Nieburg:1966:JS

[Nie66b] H. L. Nieburg. The JPL story. Bulletin of the Atomic Scientists, 22(8):35–38, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [DN67].

Nielsen:2014:WWR

[Nie14] Jenny Nielsen. Women and weapons: Redressing the gender gap: a Danish response. Bulletin of the Atomic Scientists, 70(5):17– 20, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Niles:2014:HCW

[Nil14] Ken Niles. The Hanford cleanup: What’s taking so long? Bulletin of the Atomic Scientists, 70(4):37–48, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nininger:1959:WUS

[Nin59] Robert Nininger. World uranium supplies. Bulletin of the Atomic Scientists, 15(6):249, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ninane:1988:LOC

[Nin88] G. Ninane. Letter: Old, cold war. Bulletin of the Atomic Scientists, 44(1):54, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pat87].

Nishina:1947:JSD

[Nis47] Yoshio Nishina. A Japanese scientist describes the destruction of his cyclotrons. Bulletin of the Atomic Scientists, 3(6):145, 167, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nitze:1957:AMU

[Nit57] Paul H. Nitze. Aims and methods of United States foreign policy. Bulletin of the Atomic Scientists, 13(8):292–294, 312, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nitze:1980:CNL

[Nit80] Paul H. Nitze. Commentary: ‘No longer a choice’. Bulletin of the Atomic Scientists, 36(7):6–8, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nitze:1989:ACC

[Nit89] Paul H. Nitze. Arms control criteria: Deterrence, parity, stability. Bulletin of the Atomic Scientists, 45(4):37–38, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nitze:1992:KNI

[Nit92] P. H. Nitze. Keep nuclear insurance. Bulletin of the Atomic Scientists, 48(4):34–36, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nixon:1960:D

[Nix60a] Richard M. Nixon. On disarmament. Bulletin of the Atomic Scientists, 16(6):216, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nixon:1960:SR

[Nix60b] Richard M. Nixon. The scientific revolution. Bulletin of the Atomic Scientists, 16(9):348–351, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nixdorff:2010:ATD

[Nix10] Kathryn Nixdorff. Advances in targeted delivery and the future of bioweapons. Bulletin of the Atomic Scientists, 66(1):24–33, January 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
1/24.

Nagler:1986:LSH

[NK86a] Michael N. Nagler and Joseph C. Kunkel. Letter: Sanctifying Hiroshima. Bulletin of the Atomic Scientists, 42(2):47, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei85a].

Nunn:1986:SDA

[NK86b] Sam Nunn and John Kerry. From Senate debate on an amendment to restrict funding for the Bigeye bomb. Bulletin of the Atomic Scientists, 42(9):29, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:2002:GNS

[NK02] Robert S. Norris and Hans M. Kristensen. Global nuclear stockpiles, 1945–2002. Bulletin of the Atomic Scientists, 58(6): 103–104, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/103.

Norris:2003:CNF

[NK03a] Robert S. Norris and Hans M. Kristensen. Chinese nuclear forces, 2003. Bulletin of the Atomic Scientists, 59(6):77–80, November/ December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
6/77.

Norris:2003:NP

[NK03b] Robert S. Norris and Hans M. Kristensen. Nuclear pursuits. Bulletin of the Atomic Scientists, 59(5):71–72, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/5/71.

Norris:2003:RNF

[NK03c] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2003. Bulletin of the Atomic Scientists, 59(4):70–72, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/4/70.

Norris:2003:UNF

[NK03d] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2003. Bulletin of the Atomic Scientists, 59(3):73–76, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/3/73.

Norris:2004:DUN

[NK04a] Robert S. Norris and Hans M. Kristensen. Dismantling U.S. nuclear warheads. Bulletin of the Atomic Scientists, 60(1):72–74, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/60/
1/72.

Norris:2004:RNF

[NK04b] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2004. Bulletin of the Atomic Scientists, 60(4):72–74, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/4/72.

Norris:2004:UNF

[NK04c] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2004. Bulletin of the Atomic Scientists, 60(3):68–70, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/60/3/68.

Norris:2004:UNR

[NK04d] Robert S. Norris and Hans M. Kristensen. U.S. nuclear reductions. Bulletin of the Atomic Scientists, 60(5):70–71, September/ October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/60/
5/70.

Norris:2004:UNW

[NK04e] Robert S. Norris and Hans M. Kristensen. U.S. nuclear weapons in Europe, 1954–2004. Bulletin of the Atomic Scientists, 60(6): 76–77, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/6/76.

Norris:2005:BNF

[NK05a] Robert S. Norris and Hans M. Kristensen. British nuclear forces, 2005. Bulletin of the Atomic Scientists, 61(6):77–79, November/ December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
6/77.

Norris:2005:FNF

[NK05b] Robert S. Norris and Hans M. Kristensen. French nuclear forces, 2005. Bulletin of the Atomic Scientists, 61(4):73–75, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/4/73.

Norris:2005:INF

[NK05c] Robert S. Norris and Hans M. Kristensen. India’s nuclear forces, 2005. Bulletin of the Atomic Scientists, 61(5):73–75, September/ October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
5/73.

Norris:2005:NKN

[NK05d] Robert S. Norris and Hans M. Kristensen. North Korea’s nuclear program, 2005. Bulletin of the Atomic Scientists, 61(3):64–67, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
3/64.

Norris:2005:RNF

[NK05e] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2005. Bulletin of the Atomic Scientists, 61(2):70–72, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/2/70.

Norris:2005:UNF

[NK05f] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2005. Bulletin of the Atomic Scientists, 61(1):73–75, January/ February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
1/73.

Norris:2006:GNS

[NK06a] Robert S. Norris and Hans M. Kristensen. Global nuclear stockpiles, 1945–2006. Bulletin of the Atomic Scientists, 62(4):64– 66, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/4/64.

Norris:2006:RNF

[NK06b] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2006. Bulletin of the Atomic Scientists, 62(2):64–67, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/2/64.

Norris:2006:UNF

[NK06c] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2006. Bulletin of the Atomic Scientists, 62(1):68–71, January/ February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/62/
1/68.

Norris:2006:UNT

[NK06d] Robert S. Norris and Hans M. Kristensen. U.S. nuclear threats: Then and now. Bulletin of the Atomic Scientists, 62(5):69–71, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/5/69.

Norris:2006:WB

[NK06e] Robert S. Norris and Hans M. Kristensen. Where the bombs are, 2006. Bulletin of the Atomic Scientists, 62(6):57–58, November/ December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/62/
6/57.

Norris:2007:INF

[NK07a] Robert S. Norris and Hans M. Kristensen. India’s nuclear forces, 2007. Bulletin of the Atomic Scientists, 63(4):74–78, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/63/4/74.

Norris:2007:NCM

[NK07b] Robert S. Norris and Hans M. Kristensen. Nuclear cruise missiles. Bulletin of the Atomic Scientists, 63(6):60–62, November/ December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
6/60.

Norris:2007:PNF

[NK07c] Robert S. Norris and Hans M. Kristensen. Pakistan’s nuclear forces, 2007. Bulletin of the Atomic Scientists, 63(3):71–74, May/ June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/63/3/71.

Norris:2007:RNF

[NK07d] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2007. Bulletin of the Atomic Scientists, 63(2):??, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/63/2/61.

Norris:2007:UNF

[NK07e] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2007. Bulletin of the Atomic Scientists, 63(1):79–82, January/ February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
1/79.

Norris:2007:UNS

[NK07f] Robert S. Norris and Hans M. Kristensen. The U.S. nuclear stockpile, today and tomorrow. Bulletin of the Atomic Scientists, 63(5):??, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/5/60.

Norris:2008:CNF

[NK08a] Robert S. Norris and Hans M. Kristensen. Chinese nuclear forces, 2008. Bulletin of the Atomic Scientists, 64(3):42–44, July/August 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/3/42.

Norris:2008:FNF

[NK08b] Robert S. Norris and Hans M. Kristensen. French nuclear forces, 2008. Bulletin of the Atomic Scientists, 64(4):??, September 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/4/52.

Norris:2008:INF

[NK08c] Robert S. Norris and Hans M. Kristensen. Indian nuclear forces, 2008. Bulletin of the Atomic Scientists, 64(5):38–40, November/ December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/64/
5/38.

Norris:2008:RNF

[NK08d] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2008. Bulletin of the Atomic Scientists, 64(2):54–57, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/2/54.

Norris:2008:UNF

[NK08e] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2008. Bulletin of the Atomic Scientists, 64(1):50–53, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/1/50.

Norris:2009:PNF

[NK09a] Robert S. Norris and Hans Kristensen. Pakistani nuclear forces, 2009. Bulletin of the Atomic Scientists, 65(5):82–89, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/5/82.

Norris:2009:NNW

[NK09b] Robert S. Norris and Hans M. Kristensen. Nuclear notebook: Worldwide deployments of nuclear weapons, 2009. Bulletin of the Atomic Scientists, 65(6):86–98, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/6/86.

Norris:2009:NUS

[NK09c] Robert S. Norris and Hans M. Kristensen. Nuclear U.S. and Soviet/Russian intercontinental ballistic missiles, 1959–2008. Bulletin of the Atomic Scientists, 65(1):62–69, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/1/62.

Norris:2009:RNF

[NK09d] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2009. Bulletin of the Atomic Scientists, 65(3):55–64, May 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/3/55.

Norris:2009:UNF

[NK09e] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2009. Bulletin of the Atomic Scientists, 65(2):59–69, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/2/59.

Norris:2009:UNW

[NK09f] Robert S. Norris and Hans M. Kristensen. U.S. nuclear warheads, 1945–2009. Bulletin of the Atomic Scientists, 65(4):72–81, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/4/72.

Norris:2010:CNF

[NK10a] Robert S. Norris and Hans M. Kristensen. Chinese nuclear forces, 2010. Bulletin of the Atomic Scientists, 66(6):134–141, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/6/134.

Norris:2010:GNW

[NK10b] Robert S. Norris and Hans M. Kristensen. Global nuclear weapons inventories, 1945–2010. Bulletin of the Atomic Scientists, 66(4):77–83, July/August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/66/4/77.

Norris:2010:INF

[NK10c] Robert S. Norris and Hans M. Kristensen. Indian nuclear forces, 2010. Bulletin of the Atomic Scientists, 66(5):76–81, September/ October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
5/76.

Norris:2010:RNF

[NK10d] Robert S. Norris and Hans M. Kristensen. Russian nuclear forces, 2010. Bulletin of the Atomic Scientists, 66(1):74–81, January 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/1/74.

Norris:2010:UNF

[NK10e] Robert S. Norris and Hans M. Kristensen. U.S. nuclear forces, 2010. Bulletin of the Atomic Scientists, 66(3):57–71, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/3/57.

Norris:2011:UTN

[NK11] Robert S. Norris and Hans M. Kristensen. US tactical nuclear weapons in Europe, 2011. Bulletin of the Atomic Scientists, 67(1): 64–73, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/67/1/64.

Norris:2012:CMC

[NK12a] Robert S. Norris and Hans M. Kristensen. The Cuban Missile Crisis: a nuclear order of battle, October and November 1962. Bulletin of the Atomic Scientists, 68(6):85–91, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/6/85.

Norris:2012:NP

[NK12b] Robert S. Norris and Hans M. Kristensen. Nuclear pursuits, 2012. Bulletin of the Atomic Scientists, 68(1):94–98, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/1/94.

Norris:2013:BNS

[NK13] Robert S. Norris and Hans M. Kristensen. The British nuclear stockpile, 1953–2013. Bulletin of the Atomic Scientists, 69(4):69– 75, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/69/
4/69.

Norris:2015:CNW

[NK15] Robert S. Norris and Hans M. Kristensen. Counting nuclear warheads in the public interest. Bulletin of the Atomic Scientists, 71(1):85–90, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:2016:DUN

[NK16] Robert S. Norris and Hans M. Kristensen. Declassified: US nuclear weapons at sea during the Cold War. Bulletin of the Atomic Scientists, 72(1):58–61, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nixon:2020:SPP

[NK20] Richard M. Nixon and John F. Kennedy. 1960: Science and party politics. Bulletin of the Atomic Scientists, 76(6):366–370, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1847484.

Norris:2003:BFB

[NKH03a] Robert S. Norris, Hans M. Kristensen, and Joshua Handler. The B61 family of bombs. Bulletin of the Atomic Scientists, 59(1):74– 76, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/1/74.

Norris:2003:NKN

[NKH03b] Robert S. Norris, Hans M. Kristensen, and Joshua Handler. North Korea’s nuclear program, 2003. Bulletin of the Atomic Scientists, 59(2):74–77, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/2/74.

Novick:1961:CB

[NL61] A. Novick and J. Lederberg. Challenges to biology. Bulletin of the Atomic Scientists, 17(5–6):203–206, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nijsse:2025:HLP

[NLS25] Femke J. M. M. Nijsse, Timothy M. Lenton, and Steven R. Smith. How to leverage positive tipping points for climate action. Bulletin of the Atomic Scientists, 81(2):101–106, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2025.2464437.

Newman:1949:SI

[NM49] James R. Newman and Byron S. Miller. The socialist island. Bulletin of the Atomic Scientists, 5(1):13–15, January 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the introductory chapter of the authors’ book The Control of Atomic Energy, 1948, McGraw-Hill.

Nigh:1980:TR

[NN80] Ronald B. Nigh and James D. Nations. Tropical rainforests. Bulletin of the Atomic Scientists, 36(3):12–19, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nolan:1990:MMS

[Nol90] Janne E. Nolan. Missile mania: Some rules for the game. Bulletin of the Atomic Scientists, 46(4):27–29, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Eis90].

Nolan:2017:CCM

[Nol17] Janne E. Nolan. Cold combat: The memoir of a nuclear convert. Bulletin of the Atomic Scientists, 73(3):192–195, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1315042.

Nordheim:1947:COR

[Nor47] L. W. Nordheim. Comment from Oak Ridge. Bulletin of the Atomic Scientists, 3(4–5):124, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nordheim:1954:FI

[Nor54] L. W. Nordheim. Fear and information. Bulletin of the Atomic Scientists, 10(9):342–346, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nordheim:1955:TNW

[Nor55] L. W. Nordheim. Tests of nuclear weapons. Bulletin of the Atomic Scientists, 11(7):253–255, 272, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nordau:1956:LEF

[Nor56] Claudy-Gabrielle Nordau. Letter to the Editor: France’s contribution to the Geneva Conference. Bulletin of the Atomic Scientists, 12(2):64, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nordness:1963:AFY

[Nor63] Nedville E. Nordness. ACDA: The first year of research. Bulletin of the Atomic Scientists, 19(1):36–37, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

North:1964:BRB

[Nor64] Robert C. North. Books: The race between destruction and adaptability: Man and Society in Disaster, edited by George W. Baker and Dwight W. Chapman; Behavioral Science and Civil Defense, edited by George W. Baker and Leonard S. Cottrell; The Behavioral Sciences Today, edited by Bernard Berelson; Human Behavior: An Inventory of Scientific Findings, by Bernard Berelson and Gary A. Steiner; The Scientific Renaissance, 1450– 1630, by Marie Boas; The Study of Society: A Unified Approach, by Alfred Kuhn. Bulletin of the Atomic Scientists, 20(3):26–28, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nortman:1974:QVQ

[Nor74] Dorothy Nortman. Quantity vs. quality of life: The changing economic implications of population growth. Bulletin of the Atomic Scientists, 30(6):13–16, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norman:1981:CTB

[Nor81] Liane Ellison Norman. Commentary: Not too busy. Bulletin of the Atomic Scientists, 37(3):61–62, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Eli81].

Norman:1982:CB

[Nor82] Liane Ellison Norman. Commentary: Not on any beach. Bulletin of the Atomic Scientists, 38(2):47, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mac81c].

Norman:1988:BRB

[Nor88] Colin Norman. Book review: Selling science: How the Press Covers Science and Technology, by Dorothy Nelkin. Bulletin of the Atomic Scientists, 44(4):49–50, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1994:NNEa

[Nor94a] Robert S. Norris. Nuclear notebook: Estimated Russian (C.I.S.) nuclear stockpile, September 1994. Bulletin of the Atomic Scientists, 50(5):61–63, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

North:1994:LBS

[Nor94b] Amelia P. North. Letter: Bikini — still lying after all these years. Bulletin of the Atomic Scientists, 50(4):58–59, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei94c].

Norris:2002:RSB

[Nor02a] Robert S. Norris. The rest of the story: [book review:] Russian Strategic Nuclear Forces by Paul Podvig, ed. (Cambridge, Massachusetts: MIT Press, 2001). Bulletin of the Atomic Scientists, 58(2):63–64, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/2/63;   http://
docs.nrdc.org/nuclear/nuc_02030001a_023.pdf.

Norris:2002:RBG

[Nor02b] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Steerforth Press, South Royalton, VT, USA, 2002. ISBN 1-58642-039-9. xxi + 722 pp. LCCN UG128.G76 N67 2002. URL http:

//www.loc.gov/catdir/toc/fy036/2001057629.html.

Norris:2003:RBG

[Nor03] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Turnaround, London, UK, 2003. ISBN 1-58642-067-4. 228 pp. LCCN UG128.G76 N67 2002. URL http://www.loc.gov/

catdir/toc/fy036/2001057629.html.

Norris:2004:DBR

[Nor04] Robert S. Norris. Details: [book review:] Atom Bombs: The Top Secret Inside Story of Little Boy and Fat Man by John Coster-Mullen (self-published, 2004). Bulletin of the Atomic Scientists, 60(6):74–75, November/December 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/6/74;   http://
docs.nrdc.org/nuclear/nuc_04110001a_024.pdf.

Norris:2006:GMB

[Nor06] Robert S. Norris. The godfather of molecular biology: Book review: J. D. Bernal: The Sage of Science, by Andrew Brown, Oxford University Press, 562 pages, 2005, $34.95. Bulletin of the Atomic Scientists, 62(4):59–61, July/August 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/4/59;   http://
docs.nrdc.org/nuclear/nuc_0607_bernal.pdf.

Nordhaus:2011:ACE

[Nor11] William D. Nordhaus. The architecture of climate economics: Designing a global agreement on global warming. Bulletin of the Atomic Scientists, 67(1):9–18, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/67/1/9.

Notestein:1951:NWP

[Not51] Frank W. Notestein. The needs of world population. Bulletin of the Atomic Scientists, 7(4):99–101, 128, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Notter:2015:SCB

[Not15] Dominic A. Notter. Small country, big challenge: Switzerland’s upcoming transition to sustainable energy. Bulletin of the Atomic Scientists, 71(4):51–63, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Novey:1956:FCA

[Nov56] T. Novey. FAS Congressional activity in 1955. Bulletin of the Atomic Scientists, 12(2):45, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Novick:1976:LD

[Nov76] Sheldon Novick. Letter: A disclaimer. Bulletin of the Atomic Scientists, 32(3):4, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vHWD76].

Novick:1977:LRD

[Nov77a] Richard P. Novick. Letter: Recombinant DNA research. Bulletin of the Atomic Scientists, 33(7):2, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nov77b].

Novick:1977:PCJ

[Nov77b] Richard P. Novick. Present controls are just a start: The public will not be protected until the NIH guidelines are strengthened, tightened and made universal. Bulletin of the Atomic Scientists, 33(5):16, 18–22, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Nov77a].

Noyes:1946:BAE

[Noy46a] W. Albert Noyes, Jr. The British Atomic Energy Act. Bulletin of the Atomic Scientists, 2(11–12):8–10, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Noyes:1946:UNE

[Noy46b] W. Albert Noyes, Jr. The United Nations Educational, Scientific and Cultural Organization. Bulletin of the Atomic Scientists, 2 (3–4):16–17, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Noyes:1947:UHF

[Noy47] W. Albert Noyes, Jr. UNESCO holds first general session in Paris: Report of the UNESCO conference in Paris, and the Natural Science Subcommission’s plan. Bulletin of the Atomic Scientists, 3(3):92–93, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Noyes:1957:DWN

[Noy57] W. Albert Noyes, Jr. Do we need a foreign policy in science? Bulletin of the Atomic Scientists, 13(7):234–237, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelkin:1980:FGC

[NP80] Dorothy Nelkin and Michael Pollak. French and German courts on nuclear power. Bulletin of the Atomic Scientists, 36(5):36–42, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nassauer:1990:AWN

[NPS90] Otfried Nassauer, Daniel Plesch, and David Shorr. Allies walk on nuclear eggshells. Bulletin of the Atomic Scientists, 46(6):10– 11, 46, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nelkin:1979:DED

[NR79] Dorothy Nelkin and Arie Rip. Distributing expertise: a Dutch experiment in public-interest science. Bulletin of the Atomic Scientists, 35(5):20–23, 54, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nagler:1985:LNW

[NR85] Michael N. Nagler and Dick Ringler. Letter: Nuclear war education. Bulletin of the Atomic Scientists, 41(3):61, March 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rin84c].

NWCSPC:1995:FTD

[Nuc95] Nuclear Weapons Cost Study Project Committee. Four trillion dollars and counting. Bulletin of the Atomic Scientists, 51(6):

32–52, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LaF96].

Nuccitelli:2018:HMH

[Nuc18] Dana Nuccitelli. How much and how fast will global sea level rise? Bulletin of the Atomic Scientists, 74(3):139–141, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1461894.

Nuccitelli:2020:HWK

[Nuc20] Dana Nuccitelli. How we know the Earth is warming and humans are responsible. Bulletin of the Atomic Scientists, 76(3): 140–144, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1751969.

Nussbaum:1979:RJ

[Nus79] Rudi H. Nussbaum. Responsible journalism. Bulletin of the Atomic Scientists, 35(7):63–64, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Fel79l, Fai79].

Nussbaum:1980:CSM

[Nus80] Rudi H. Nussbaum. Commentary: Stab misdirected. Bulletin of the Atomic Scientists, 36(8):62, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fai79].

Nussbaum:1985:SSR

[Nus85] Rudi H. Nussbaum. Survivor studies and radiation standards. Bulletin of the Atomic Scientists, 41(7):68–71, August 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nilsson:1970:CSD

[NW70]  Agnar Nilsson and Gunnar Walinder.  Communications:

Strontium-90 dosages and infant mortality … an exchange. Bulletin of the Atomic Scientists, 26(5):40, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69c].

Neuwirth:1979:CAR

[NW79] Gernot Neuwirth and Peter Weish. Commentary from Austria: 1978 referendum. Bulletin of the Atomic Scientists, 35(5):49, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pat79b].

Nye:1977:TPN

[Nye77] Joseph S. Nye. Time to plan for the next generation of nuclear technology. Bulletin of the Atomic Scientists, 33(8):38–41, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nyerere:1979:TTU

[Nye79] Julius K. Nyerere. From Tanzania: A trade union for the poor. Bulletin of the Atomic Scientists, 35(6):38–39, June 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nye:1982:NPP

[Nye82] Joseph S. Nye, Jr. Nuclear proliferation in the 1980s: Political solutions. Bulletin of the Atomic Scientists, 38(7):30–32, August/ September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nye:1989:ACC

[Nye89] Joseph S. Nye, Jr. Arms control criteria: Don’t count on counting. Bulletin of the Atomic Scientists, 45(4):43, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Nye:2013:BBC

[Nye13] Joseph S. Nye, Jr. From bombs to bytes: Can our nuclear history inform our cyber future? Bulletin of the Atomic Scientists, 69 (5):8–14, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/5/8.

OBrien:1948:LTG

[O’B48] John Lord O’Brien. Loyalty tests and guilt by association. Bulletin of the Atomic Scientists, 4(6):166–172, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OBrien:1969:PTC

[O’B69a] William O’Brien. Pharmaceutical testing: Critic’s view: The way it really is. Bulletin of the Atomic Scientists, 25(6):13–16, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OBrien:1969:DTT

[O’B69b] William M. O’Brien. Drug testing: Is time running out? Bulletin of the Atomic Scientists, 25(1):8–14, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OBrien:1975:BBF

[O’B75] R. D. O’Brien. Books: Genetic Fix, by Amitai Etzioni. Bulletin of the Atomic Scientists, 31(5):48–49, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Owens:1978:LHG

[OB78] Gwinn Owens and Stanley A. Blumberg. Letter: How got the bomb? Bulletin of the Atomic Scientists, 34(9):62, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber78a].

OBrien:1979:CRN

[O’B79] John N. O’Brien. Commentary: Regulating nuclear security. Bulletin of the Atomic Scientists, 35(9):57–58, November 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OBrien:1980:CBC

[O’B80] John N. O’Brien. Commentary: The Progressive case. Bulletin of the Atomic Scientists, 36(8):60–61, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TR80].

OBrien:1982:CFI

[O’B82] John N. O’Brien. Commentary: Fuel without ideology. Bulletin of the Atomic Scientists, 38(4):51, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Obasi:1985:AUD

[Oba85] G. O. P. Obasi. Africa: Understanding the drought. Bulletin of the Atomic Scientists, 41(8):43–45, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1949:AAC

[OBC^{+}49] J. Robert Oppenheimer, Oliver E. Buckley, James B. Conant, Lee A. DuBridge, Enrico Fermi, I. I. Rabi, Hartley Rowe, Glenn T. Seaborg, and Cyril S. Smith. AEC Advisory Committee statement on fellowships. Bulletin of the Atomic Scientists, 5 (8–9):210, 254, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Obert:1974:CGW

[Obe74] Kjell Obert. Communications: Guest workers in Sweden. Bulletin of the Atomic Scientists, 30(1):48, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gre73a].

Oberoi:1997:KB

[Obe97a] Surinder Singh Oberoi. Kashmir is bleeding. Bulletin of the Atomic Scientists, 53(2):24–32, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fai97].

Oberoi:1997:TBI

[Obe97b] Surinder Singh Oberoi. Three big ifs. Bulletin of the Atomic Scientists, 53(2):29, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oberoi:1998:RKC

[Obe98] Surinder Singh Oberoi. Report: Kashmir: Caught in the crossfire. Bulletin of the Atomic Scientists, 54(5):13–15, September/ October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Obukhov:1989:PCS

[Obu89] Alexei Obukhov. Perspective: Chief Soviet negotiator: Let’s get back to the table. Bulletin of the Atomic Scientists, 45(3):6, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OConnell:1987:BRB

[O’C87] Mary O’Connell. Book review: The Whale and the Reactor: A Search for Limits in an Age of High Technology, by Langdon Winner. Bulletin of the Atomic Scientists, 43(2):57–58, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oden:1988:MSE

[Ode88] Michael Dee Oden. Military spending erodes real national security. Bulletin of the Atomic Scientists, 44(5):36–41, June 1988.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Odhiambo:1979:KHB

[Odh79] Thomas R. Odhiambo. From Kenya: How to bridge the scientific gap. Bulletin of the Atomic Scientists, 35(6):35–38, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Odhiambo:1981:UNU

[Odh81] Thomas R. Odhiambo. Use and non-use of insects. Bulletin of the Atomic Scientists, 37(7):31–36, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oelrich:2012:NSA

[Oel12] Ivan Oelrich. The next step in arms control: Eliminate the counterforce mission. Bulletin of the Atomic Scientists, 68(1):79– 85, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/1/79.

Oelrich:2014:IFV

[Oel14] Ivan Oelrich. The Insurer’s Fallacy and the value of nuclear weapons. Bulletin of the Atomic Scientists, 70(4):69–73, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oelrich:2020:CYJ

[Oel20] Ivan Oelrich. Cool your jets: Some perspective on the hyping of hypersonic weapons. Bulletin of the Atomic Scientists, 76(1):37–45, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1701283.

Offner:1984:CWL

[Off84] Franklin F. Offner. Commentary: When less is more. Bulletin of the Atomic Scientists, 40(8):47, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Offner:2002:ASV

[Off02] Arnold A. Offner. Another such victory: President Truman and the Cold War, 1945–1953. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 2002. ISBN 0-8047-4254-

5. xv + 626 pp. LCCN E813 .O36 2002. URL http://www.

loc.gov/catdir/enhancements/fy0710/2001049130-b.html;
 http://www.loc.gov/catdir/enhancements/fy0710/2001049130-
d.html.

Oftedal:1971:CCT

[Oft71] Per Oftedal. Communications: Cancer threshold. Bulletin of the Atomic Scientists, 27(1):52, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro70].

Ouagrham-Gormley:2010:BCM

[OG10] Sonia Ben Ouagrham-Gormley. Breaking out of the CTR mold? Bulletin of the Atomic Scientists, 66(1):12–17, January 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/1/12.

Ogawa:1958:RMN

[Oga58] Iwao Ogawa. Radiation and man: Nuclear tests and radiation: Fall-out and rice contamination in Japan. Bulletin of the Atomic Scientists, 14(1):35–38, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ogawa:1968:NSC

[Oga68] Iwao Ogawa. Nuclear submarines: Comment from Japan. Bulletin of the Atomic Scientists, 24(4):37–38, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ogburn:1948:SSC

[Ogb48] William F. Ogburn. A social scientist comments. Bulletin of the Atomic Scientists, 4(3):74–75, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bri48a].

Ouagrham-Gormley:2018:BSS

[OGFM18] Sonia Ben Ouagrham-Gormley and Shannon Fye-Marnien. The bright side of synthetic biology and Crispr. Bulletin of the Atomic Scientists, 74(1):19–26, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1413056.

Oberg:1984:CSW

[OGP84] James E. Oberg, Richard Garwin, and John Pike. Commentary: Space weapons. Bulletin of the Atomic Scientists, 40(7):

47–50, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GP84, Vel84].

Ordway:1985:LNC

[OH85] Frederick I. Ordway, III and Linda Hunt. Letter: Nazi coverup. Bulletin of the Atomic Scientists, 41(11):66–67, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hun85b].

OHanlon:2001:PNN

[O’H01] Michael O’Hanlon. Play nice with NATO. Bulletin of the Atomic Scientists, 57(1):37–38, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/37.

Oishi:1972:GWW

[Ois72] Buichi Oishi. GNP — for whom? For what? Bulletin of the Atomic Scientists, 28(7):51–52, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliphant:1947:SSA

[OJCP^{+}47] M. L. Oliphant, F. Joliot-Curie, R. E. Peierls, Albert Einstein, Federation of American Scientists, and General Leslie Groves. Statements on the second anniversary of Hiroshima. Bulletin of the Atomic Scientists, 3(9):235–236, 252, September 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OKeefe:1969:MLT

[O’K69] John A. O’Keefe. Manned landings and theories of lunar formation. Bulletin of the Atomic Scientists, 25(7):56–58, 60, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OKeefe:1973:AAF

[O’K73] John A. O’Keefe. After Apollo: Fission origin of the Moon. Bulletin of the Atomic Scientists, 29(9):26–29, November 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OKeefe:1984:IWP

[O’K84] Timothy J. O’Keefe. Institution-wdie programs. Bulletin of the Atomic Scientists, 40(10):23s–24s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Okoye:1982:HA

[Oko82] S. E. Okoye. The hunger alert. Bulletin of the Atomic Scientists, 38(1):28–29, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OLeary:1973:SSN

[O’L73] Brian O’Leary. The Space Shuttle: NASA’s white elephant in the sky. Bulletin of the Atomic Scientists, 29(2):36–43, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OLeary:1975:RTE

[O’L75] Brian O’Leary. R&D: The thin end of the wedge. Bulletin of the Atomic Scientists, 31(8):8–14, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Abr76].

OLeary:1977:PC

[O’L77] Brian O’Leary. Project Columbus 1992. Bulletin of the Atomic Scientists, 33(3):4–5, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev77c].

Olanoff:1982:CPS

[Ola82] Samuel Olanoff. Commentary: Power to say no. Bulletin of the Atomic Scientists, 38(1):48, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oldendorf:1962:ADW

[Old62] W. H. Oldendorf. On the acceptability of a device as a weapon. Bulletin of the Atomic Scientists, 18(1):35–37, January 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oldham:1966:SE

[Old66] C. H. G. Oldham. Science and education. Bulletin of the Atomic Scientists, 22(6):41–50, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oldham:1969:STM

[Old69] C. H. G. Oldham. Science travels the Mao road. Bulletin of the Atomic Scientists, 25(2):80–83, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Oldershaw:1981:CGL

[Old81] Robert L. Oldershaw. Commentary: Gravity and levity. Bulletin of the Atomic Scientists, 37(7):53–54, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliphant:1948:WCM

[Oli48] M. L. Oliphant. Western control means Eastern supremacy. Bulletin of the Atomic Scientists, 4(12):365, December 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fri48a].

Oliphant:1952:SBE

[Oli52] M. L. Oliphant. Some British experiences. Bulletin of the Atomic Scientists, 8(7):230, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Oliphant:1958:CPE

[Oli58] Marcus L. Oliphant. Clever politics or emotional insight? Bulletin of the Atomic Scientists, 14(6):210–211, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliphant:1966:RPT

[Oli66] Mark Oliphant. Reports: Over pots of tea: Excerpts from a diary of a visit to China. Bulletin of the Atomic Scientists, 22(5):36–43, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oliphant:1982:BCN

[Oli82] Mark Oliphant. The beginning: Chadwick and the neutron. Bulletin of the Atomic Scientists, 38(10):14–18, December 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Par83].

Oliphant:1989:FHT

[Oli89] Mark Oliphant. Footnote to history: Three men and the bomb. Bulletin of the Atomic Scientists, 45(2):41–42, March 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olken:1977:LNB

[Olk77] Hyman Olken. Letter: The neutron bomb. Bulletin of the Atomic Scientists, 33(9):2, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olson:1959:BBP

[Ols59] Elder Olson. Books: The Poet’s Way of Knowledge, by C. Day Lewis. Bulletin of the Atomic Scientists, 15(2):92–93, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olson:1962:BBM

[Ols62a] James C. Olson. Books: From the Marco Polo Bridge to Pearl Harbor, by David J. Lu. Bulletin of the Atomic Scientists, 18(9): 28, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olson:1962:BBS

[Ols62b] James C. Olson. Books: Japan Subdued: The Atomic Bomb and the End of the War in the Pacific, by Herbert Freis. Bulletin of the Atomic Scientists, 18(2):34–35, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Opp62b] and response [Ols62c].

Olson:1962:LEJ

[Ols62c] James C. Olson. Letter to the Editor: Japan subdued. Bulletin of the Atomic Scientists, 18(6):43, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olson:1963:REA

[Ols63] Warren E. Olson. Responsibility: An escape and an approach. Bulletin of the Atomic Scientists, 19(3):2–6, March 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response and reply [HFO63].

Olsen:1974:ECB

[Ols74a] Leif H. Olsen. The energy crisis and the balance of payments. Bulletin of the Atomic Scientists, 30(3):26–29, March 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dav74, She74a].

Olson:1974:LM

[Ols74b] Robert L. Olson. Letters: A ‘Model T’. Bulletin of the Atomic Scientists, 30(6):5–6, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olson:2016:YCC

[Ols16] Julia Olson. Youth and climate change: an advocate’s argument for holding the US government’s feet to the fire. Bulletin of the Atomic Scientists, 72(2):79–84, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oltcher:1952:LEB

[Olt52] Olive Oltcher. Letter to the Editor: [Biology and the burden of our times]. Bulletin of the Atomic Scientists, 8(8):293, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ros52].

Oluwasanmi:1966:SPA

[Olu66] H. A. Oluwasanmi. On the social problems of agricultural change. Bulletin of the Atomic Scientists, 22(9):21, November 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olum:1971:BBA

[Olu71] Paul Olum. Books: Rebels Against War: The American Peace Movement 1941–1960, by Lawrence S. Wittner. Bulletin of the Atomic Scientists, 27(3):47–48, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ortmeyer:1997:WTW

[OM97] Pat Ortmeyer and Arjun Makhijani. Worse than we knew. Bulletin of the Atomic Scientists, 53(6):46–50, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ONeill:1989:PCH

[O’N89a] Dan O’Neill. Project Chariot: how Alaska escaped nuclear excavation. Bulletin of the Atomic Scientists, 45(10):28–37, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Han90d].

ONeill:1989:SM

[O’N89b] Dan O’Neill. Shooting the moon. Bulletin of the Atomic Scientists, 45(10):31, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ONeill:1994:FB

[O’N94] Dan (Daniel T.) O’Neill. The Firecracker Boys. St. Martin’s Press, New York, NY, USA, 1994. ISBN 0-312-11183-5, 0-312-11086-3. 388 pp. LCCN UG1282.A8 O39 1994.

ONeill:1995:BZT

[O’N95] Kevin O’Neill. Bulletin: On the zirconium trail. Bulletin of the Atomic Scientists, 51(5):6–7, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ooms:1946:AEUa

[Oom46a] Casper Ooms. Atomic energy and U.S. patent policy: Part 1: History of the patent system. Bulletin of the Atomic Scientists, 2 (9–10):28–29, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ooms:1946:AEUb

[Oom46b] Casper W. Ooms. Atomic energy and U.S. patent policy: Part 2: Patent provisions of the Atomic Energy Act. Bulletin of the Atomic Scientists, 2(11–12):30–31, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orlov:1998:MSG

[OP98] Vladim Orlov and William C. Potter. The mystery of the sunken gyros. Bulletin of the Atomic Scientists, 54(6):34–39, November/ December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ophuls:1973:RL

[Oph73] William Ophuls. The return of Leviathan. Bulletin of the Atomic Scientists, 29(3):50–52, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1946:ICAa

[Opp46] J. Robert Oppenheimer. The international control of atomic energy. Bulletin of the Atomic Scientists, 1(12):1–5, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Opp63c].

Oppenheimer:1947:COT

[Opp47a] Ernest Oppenheimer. The challenge of our time. Bulletin of the Atomic Scientists, 3(12):370–371, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rab47c].

Oppenheimer:1947:FIAb

[Opp47b] J. Robert Oppenheimer. Functions of the International Agency in research and development. Bulletin of the Atomic Scientists, 3 (7):173–176, 197, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1948:ICAb

[Opp48a] Robert Oppenheimer. International control of atomic energy. Bulletin of the Atomic Scientists, 4(2):39–43, 48, February 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Opp63b].

Oppenheimer:1948:PCW

[Opp48b] Robert Oppenheimer. Physics in the contemporary world. Bulletin of the Atomic Scientists, 4(3):65–68, 85–86, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1949:LSM

[Opp49a] J. R. Oppenheimer. A letter to Senator McMahon. Bulletin of the Atomic Scientists, 5(6–7):163, 178, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1949:OM

[Opp49b] J. Robert Oppenheimer. The open mind. Bulletin of the Atomic Scientists, 5(1):3–5, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See minor erratum [Ano49o].

Oppenheimer:1950:FD

[Opp50] J. Robert Oppenheimer. Fateful decision. Bulletin of the Atomic Scientists, 6(3):74, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1951:CMV

[Opp51a] J. Robert Oppenheimer. Comments on the military value of the atom. Bulletin of the Atomic Scientists, 7(2):43–45, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1951:DOR

[Opp51b] J. Robert Oppenheimer. Dr. Oppenheimer on the rate of American bomb production. Bulletin of the Atomic Scientists, 7(2): 45, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1951:SPA

[Opp51c] J. Robert Oppenheimer. The scientist in public affairs: Encouragement of science. Bulletin of the Atomic Scientists, 7(1):6–8, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Address delivered at the Awards Banquet of the Science Talent Institute in Washington, DC, March 6, 1950.

Oppenheimer:1953:AWAb

[Opp53] J. Robert Oppenheimer. Atomic weapons and American policy. Bulletin of the Atomic Scientists, 9(6):202–205, July 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Opp63a].

Oppenheimer:1954:RAV

[Opp54a] J. Robert Oppenheimer. On receiving the AEC verdict. Bulletin of the Atomic Scientists, 10(7):258, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1954:OR

[Opp54b] J. Robert Oppenheimer. Oppenheimer replies. Bulletin of the Atomic Scientists, 10(5):177–187, 191, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See charges [Nic54b].

Oppenheimer:1955:PASa

[Opp55] J. Robert Oppenheimer. Prospects in the arts and sciences. Bulletin of the Atomic Scientists, 11(2):42–44, 52, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Talk broadcast for the Columbia University Bicentennial, 26 December 1954.

Oppenheimer:1956:SOTb

[Opp56] J. Robert Oppenheimer. Science and our times. Bulletin of the Atomic Scientists, 12(7):235–237, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Talk delivered to the Roosevelt University Founders and Friends Dinner in Chicago, 22 May 1956.

Oppenheimer:1959:GSS

[Opp59a] J. Robert Oppenheimer. The growth of science and the structure of culture. Bulletin of the Atomic Scientists, 15(2):76, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Two-paragraph excerpt from the author’s article of the same name in Daedalus, winter 1958.

Oppenheimer:1959:SVH

[Opp59b] J. Robert Oppenheimer. Science, values, and the human community. Bulletin of the Atomic Scientists, 15(1):7, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1960:KU

[Opp60] J. Robert Oppenheimer. In the keeping of unreason. Bulletin of the Atomic Scientists, 16(1):18–22, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This article was given at a seminar on “Industrial Society and the Western Political Dialogue”, help by the Congress for Cultural Freedom at Rheinfelden, Switzerland, September 20–27, 1959.

Oppenheim:1962:BBS

[Opp62a] A. L. Oppenheim. Books: Science Since Babylon, by Derek J. de Solla Price. Bulletin of the Atomic Scientists, 18(3):41–42, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1962:LEJ

[Opp62b] J. Robert Oppenheimer. Letter to the Editor: Japan subdued. Bulletin of the Atomic Scientists, 18(6):41, 43, June 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ols62b].

Oppenheimer:1963:AWA

[Opp63a] J. Robert Oppenheimer. Atomic weapons and American policy. In Grodzins and Rabinowitch [GR63], pages 188–196. LCCN D842 .B78. Reprint of [Opp53].

Oppenheimer:1963:FIC

[Opp63b] J. Robert Oppenheimer. The failure of international control. In Grodzins and Rabinowitch [GR63], pages 64–75. LCCN D842 .B78. Reprint of [Opp48a].

Oppenheimer:1963:ICA

[Opp63c] J. Robert Oppenheimer. The international control of atomic energy. In Grodzins and Rabinowitch [GR63], pages 53–63. LCCN D842 .B78. Reprint of [Opp46].

Oppenheimer:1963:NB

[Opp63d] J. Robert Oppenheimer. Niels Bohr. Bulletin of the Atomic Scientists, 19(7):9, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1963:TC

[Opp63e] J. Robert Oppenheimer. A talk in Chicago. Bulletin of the Atomic Scientists, 19(8):4–6, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). First presented as a talk, “The Scientific Revolution and Its Effects on Democratic Institutions”, in Chicago on June 13, 1963, as part of the symposium on “Challenges to Democracy”, sponsored by the Fund for the Republic.

Oppenheimer:1968:WSH

[Opp68] Frank Oppenheimer. A war in the shadow of the H-bomb. Bulletin of the Atomic Scientists, 24(5):43–45, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1979:CSB

[Opp79a] J. Robert Oppenheimer. Celebration of the sixtieth birthday of Albert Einstein. Bulletin of the Atomic Scientists, 35(3):39, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Oppenheimer:1979:OE

[Opp79b] J. Robert Oppenheimer. Oppenheimer on Einstein. Bulletin of the Atomic Scientists, 35(3):36–38, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Oppenheimer:1990:LB

[Opp90a] Andy Oppenheimer. Letter: Brilliant! Bulletin of the Atomic Scientists, 46(1):45, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:1990:LRS

[Opp90b] Peter Oppenheimer. Letter: Right from the start. Bulletin of the Atomic Scientists, 46(8):46, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oppenheimer:2002:NIC

[Opp02a] A. R. Oppenheimer. Northern Ireland: The countdown begins. Bulletin of the Atomic Scientists, 58(3):13–15, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/3/13.

Oppenheimer:2002:BTT

[Opp02b] Andy Oppenheimer. Britain: Terror target number two? Bulletin of the Atomic Scientists, 58(5):12–14, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/58/5/12.

Oppenheimer:2003:BHM

[Opp03a] Andy Oppenheimer. Britain: Her Majesty’s new nukes? Bulletin of the Atomic Scientists, 59(2):16–18, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/2/16.

Oppenheimer:2003:BSS

[Opp03b] Andy Oppenheimer. Britain: Sellafield, Salmon, and the Irish sea. Bulletin of the Atomic Scientists, 59(5):11–13, September/ October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
5/11.

Oppenheimer:2004:MNB

[Opp04] Andy Oppenheimer. Mini-nukes: Boom or bust? Bulletin of the Atomic Scientists, 60(5):12–14, September/October 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/5/12.

Oppenheimer:2020:SOT

[Opp20] J. Robert Oppenheimer. 1956: Science and our times. Bulletin of the Atomic Scientists, 76(6):356–358, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1847481.

Oppenheimer:2023:OR

[Opp23] J. R. Oppenheimer. Oppenheimer replies. Bulletin of the Atomic Scientists, 79(4):242–254, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2223084.

Owen:1970:IOS

[OR70] Tobias Owen and David L. Roberts. Inner and outer space: New horizon in space exploration. Bulletin of the Atomic Scientists, 26(2):23–29, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orans:1965:LEU

[Ora65] Martin Orans. Letter to the Editor: A U.S.–Soviet Kula. Bulletin of the Atomic Scientists, 21(3):44–45, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1958:BBN

[Ore58a] Jay Orear. Books: Our Nuclear Future. Facts, Dangers, and Opportunities, by Edward Teller. Bulletin of the Atomic Scientists, 14(6):235–236, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1958:DNW

[Ore58b] Jay Orear. Detection of nuclear weapons testing. Bulletin of the Atomic Scientists, 14(3):98–101, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Wes58, Wig58].

Orear:1959:TBT

[Ore59] Jay Orear. Test ban: Two views: How feasible is a test ban? Bulletin of the Atomic Scientists, 15(3):99–102, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1960:AGT

[Ore60a] Jay Orear. An appraisal of the Geneva talks. Bulletin of the Atomic Scientists, 16(2):50–51, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lof60].

Orear:1960:RFC

[Ore60b] Jay Orear. Report on the “Foothils” Conference — the test-ban breakthrough. Bulletin of the Atomic Scientists, 16(4):132–133, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1961:NAI

[Ore61a] Jay Orear. A new approach to inspection. Bulletin of the Atomic Scientists, 17(3):107–110, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1961:RO

[Ore61b] Jay Orear. Report from Oslo. Bulletin of the Atomic Scientists, 17(7):273–274, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1962:LEF

[Ore62a] Jay Orear. Letter to the Editor: Follow-up — re: Strontium. Bulletin of the Atomic Scientists, 18(2):32, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1962:LES

[Ore62b] Jay Orear. Letter to the Editor: Over the strontium 90 limit? Bulletin of the Atomic Scientists, 18(1):37, January 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orear:1963:SZD

[Ore63] Jay Orear. Safeguarded zonal disarmament. Bulletin of the Atomic Scientists, 19(2):18–21, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orleans:1966:DPP

[Orl66] Leo A. Orleans. Dealing with population problems. Bulletin of the Atomic Scientists, 22(6):22–26, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orleans:1983:CPC

[Orl83] Leo A. Orleans. Changing perceptions of China. Bulletin of the Atomic Scientists, 39(2):13–15, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fai83].

Orleans:1984:STC

[Orl84] Leo A. Orleans. Science and technology in China: Education, careers and social status. Bulletin of the Atomic Scientists, 40 (8):11s–13s [various between 24–25], October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ormond:1957:LED

[Orm57] Eugene Ormond. Letter to the Editor: Do as the teacher says. Bulletin of the Atomic Scientists, 13(5):185, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ormsby:1985:AM

[Orm85] David Ormsby. Abandon the MX. Bulletin of the Atomic Scientists, 41(4):64, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ornstein:1967:FPE

[Orn67] Leonard Ornstein. Forum: The population explosion, “Conservative eugenics”, and human evolution. Bulletin of the Atomic Scientists, 23(6):57–60, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Led66a, Led67].

Ornstein:1985:LSC

[Orn85] Severo M. Ornstein. Letter: Strategic computing. Bulletin of the Atomic Scientists, 41(6):56, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Der85].

Orowan:1959:OUS

[Oro59] Egon Orowan. Our universities and scientific creativity. Bulletin of the Atomic Scientists, 15(6):236–239, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orr:1956:BPG

[Orr56a] Lord Boyd Orr. The British Parliamentary Group for World Government. Bulletin of the Atomic Scientists, 12(1):21–22, 24, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orr:1956:PIS

[Orr56b] Lord Boyd Orr. Part IV: Science and world economy: Science and hunger. Bulletin of the Atomic Scientists, 12(8):308–311, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orr:1978:PNP

[Orr78] John Boyd Orr. Perspective: Nobel Peace Laureates: 1949. Bulletin of the Atomic Scientists, 34(9):1, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 12 December 1949, Oslo, Norway.

OSullivan:1963:DRI

[O’S63] Thomas C. O’Sullivan, Jr. Disadvantages of reliable inspection. Bulletin of the Atomic Scientists, 19(3):18–19, March 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Otero:1981:COU

[OS81] Mario Otero and Barbara Sureda. Commentary: On oppression in Uruguay. Bulletin of the Atomic Scientists, 37(3):63–64, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Addendum to [Ote81].

OShea:1986:RNB

[O’S86] James O’Shea. The real nuts and bolts of Pentagon contracts. Bulletin of the Atomic Scientists, 42(8):19–20, October 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Orlov:2000:NWR

[OS00] Vladimir Orlov and Ivan A. Safranchuk. Nuclear weapons: The Russian public speaks. Bulletin of the Atomic Scientists, 56(1): 16–18, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/1/16.

Osamu:1977:SSF

[Osa77] Kataoka Osamu. A survivor’s story: “Friends, please forgive us”. Bulletin of the Atomic Scientists, 33(10):52, December 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborn:1947:RDA

[Osb47] Frederick H. Osborn. The Russians delay action on atomic control. Bulletin of the Atomic Scientists, 3(10):299–300, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborn:1949:AER

[Osb49a] Frederick Osborn. The atomic explosion in Russia: The United Nations faces the new situation. Bulletin of the Atomic Scientists, 5(10):267, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborn:1949:CAE

[Osb49b] Frederick H. Osborn. Control of atomic energy: An international problem. Bulletin of the Atomic Scientists, 5(8–9):251–254, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborn:1954:GET

[Osb54] Fairfield Osborn. Guest editorial: Their need — our obligation. Bulletin of the Atomic Scientists, 10(3):73–74, March 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborne:1958:RMB

[Osb58] Louis S. Osborne. Radiation and man: Broad studies: The National Academy and Medical Research Council reports. Bulletin of the Atomic Scientists, 14(1):46–47, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborn:1982:CAA

[Osb82] Earl D. Osborn. Commentary: An argument against espionage. Bulletin of the Atomic Scientists, 38(6):69, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osborn:1983:UW

[Osb83] Earl D. Osborn. From an unconventional warrior. Bulletin of the Atomic Scientists, 39(7):7–8, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osgood:1960:CGU

[Osg60] Charles E. Osgood. A case for graduated unilateral disengagement. Bulletin of the Atomic Scientists, 16(4):127–131, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Sha61].

Osgood:1962:RDN

[Osg62] Charles E. Osgood. Rational defense: Nuclear displacement. Bulletin of the Atomic Scientists, 18(6):21–24, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osgood:1967:OCP

[Osg67] Charles E. Osgood. Our crises in perspective. Bulletin of the Atomic Scientists, 23(2):12–16, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osgood:1980:CGS

[Osg80] Charles E. Osgood. Commentary: The GRIT strategy. Bulletin of the Atomic Scientists, 36(5):58–60, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osgood:1983:EHP

[Osg83] Eugenia V. Osgood. Euromissiles: historical and political realities. Bulletin of the Atomic Scientists, 39(10):13–21, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [McA84].

Osgood:1985:HNR

[Osg85] Charles Osgood. Hiroshima and Nagasaki revisited. Bulletin of the Atomic Scientists, 41(7):63–67, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Osnos:1977:DVM

[Osn77] Peter Osnos. Détente: a victim of mutual suspicion: an American correspondent’s analysis of what went wrong in U.S.–Soviet relations. Bulletin of the Atomic Scientists, 33(9):16–19, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ornstein:1984:SC

[OSS84] Severo M. Ornstein, Brian C. Smith, and Lucy A. Suchman. Strategic computing. Bulletin of the Atomic Scientists, 40(10): 11–15, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CO85, Der85].

Ostow:1963:WU

[Ost63] Mortimer Ostow. War and the unconscious. Bulletin of the Atomic Scientists, 19(1):25–28, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ostrom:1967:PFN

[Ost67] Vincent Ostrom. Political feasibility [of Nawapa]. Bulletin of the Atomic Scientists, 23(7):13–17, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Oster:1972:BBA

[Ost72] Patrick R. Oster. Books: Science in American Society: A Social History, by George H. Daniels. Bulletin of the Atomic Scientists, 28(2):47–48, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OToole:2000:AS

[O’T00] Tara O’Toole. Alice Stewart. Bulletin of the Atomic Scientists, 56 (4):66–68, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/56/4/66.

Otero:1981:OU

[Ote81] Mario Otero. Oppression in Uruguay. Bulletin of the Atomic Scientists, 37(2):29–31, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [OS81].

Ouellette:2007:DE

[Oue07] Jennifer Ouellette. The danger equation. Bulletin of the Atomic Scientists, 63(5):17, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/5/17.

Orear:1965:ESS

[OW65] Jay Orear and Lincoln Wolfenstein. European scientists speak. Bulletin of the Atomic Scientists, 21(4):44–45, April 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olshansky:1990:RHC

[OW90a] S. Jay Olshansky and R. Gary Williams. Radiation and health: Culture shock at the weapons complex. Bulletin of the Atomic Scientists, 46(7):29–33, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Olshansky:1990:UD

[OW90b] S. Jay Olshansky and R. Gary Williams. A “universe” of data. Bulletin of the Atomic Scientists, 46(7):31, September 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

OCallaghan:1999:WD

[OW99] Geraldine O’Callaghan and Brian Wood. Wheeling and dealing. Bulletin of the Atomic Scientists, 55(1):52–54, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/1/52.

Ogilvie-White:2015:ARN

[OW15] Tanya Ogilvie-White. Australia’s rocky nuclear past and uncertain future. Bulletin of the Atomic Scientists, 71(5):59– 66, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/59.full.pdf+html.

Owen:1954:OF

[Owe54] David Owen. “Operation Faith”. Bulletin of the Atomic Scientists, 10(3):81–83, 112, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Owen:1979:TUP

[Owe79] Wilfrid Owen. Transition to an urban planet. Bulletin of the Atomic Scientists, 35(9):12–18, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Owen:1987:LST

[Owe87] John Owen. Letter: Stop testing now. Bulletin of the Atomic Scientists, 43(2):60, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol87a].

Ozersky:1992:BRT

[Oze92] Josh Ozersky. Book review: TV guide: Is Anyone Responsible?, by Shanto Iyengar. Bulletin of the Atomic Scientists, 48(7):43, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pabian:1988:RES

[Pab88] Frank Pabian. Reviewing the evidence on South Africa and the bomb. Bulletin of the Atomic Scientists, 44(3):47–48, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pachter:1959:LER

[Pac59] Henry M. Pachter. Letter to the Editor: Recognizing the Status Quo. Bulletin of the Atomic Scientists, 15(9):401, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab59d].

Page:1961:WD

[Pag61] Thornton Page. Where are the dolphins? Bulletin of the Atomic Scientists, 17(9):392–393, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hut61].

Paget:1968:SRA

[Pag68] Roger K. Paget. Special reports from Asia: Indonesian politics: The new order emerges. Bulletin of the Atomic Scientists, 24 (2):30–34, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Page:1969:VO

[Pag69] Thornton Page. A view from the outside. Bulletin of the Atomic Scientists, 25(7):61–66, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1980:PIA

[Pai80a] Christopher Paine. Pershing II: the Army’s strategic weapon. Bulletin of the Atomic Scientists, 36(8):25–31, October 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano80i].

Paine:1980:MPW

[Pai80b] Christopher E. Paine. MX: the public works project of the 1980s. Bulletin of the Atomic Scientists, 36(2):12–16, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1981:CMR

[Pai81a] Christopher Paine. Commentary: On “The Mouse that Roared”: The ‘Shuman Plan’. Bulletin of the Atomic Scientists, 37(4): 59–64, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shu81] and comment [Swe81].

Paine:1981:RCM

[Pai81b] Christopher Paine. Running in circles with the MX. Bulletin of the Atomic Scientists, 37(10):5–10, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SP82].

Paine:1982:AIT

[Pai82a] Christopher Paine. Admiral Inman’s tidal wave. Bulletin of the Atomic Scientists, 38(3):3–6, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gus82].

Paine:1982:CWB

[Pai82b] Christopher Paine. Commentary from Washington: Bomber mania. Bulletin of the Atomic Scientists, 38(1):4–5, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:EAB

[Pai82c] Christopher Paine. Election 1982: Arms buildup. Bulletin of the Atomic Scientists, 38(8):5–8, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:EMS

[Pai82d] Christopher Paine. The elusive ‘margin of safety’. Bulletin of the Atomic Scientists, 38(5):11–14, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:FS

[Pai82e] Christopher Paine. A false START. Bulletin of the Atomic Scientists, 38(7):11–14, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:FUN

[Pai82f] Christopher Paine. The freeze and the United Nations. Bulletin of the Atomic Scientists, 38(6):10–15, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:LRN

[Pai82g] Christopher Paine. Last roundup for NATO? Bulletin of the Atomic Scientists, 38(2):3–7, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:NSP

[Pai82h] Christopher Paine. National security and political freedom: Secrets. Bulletin of the Atomic Scientists, 38(4):11–16, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:NCF

[Pai82i] Christopher Paine. Nuclear combat: the five-year defense plan. Bulletin of the Atomic Scientists, 38(9):5–12, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1982:PSC

[Pai82j] Christopher Paine. Pugwash 1982: Statement from the Council. Bulletin of the Atomic Scientists, 38(9):2, November 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1983:ATL

[Pai83a] Christopher Paine. The ABM treaty: looking for loopholes. Bulletin of the Atomic Scientists, 39(7):13–16, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1983:BNP

[Pai83b] Christopher Paine. Build now, pay later. Bulletin of the Atomic Scientists, 39(9):6–9, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1983:CRE

[Pai83c] Christopher Paine. Curtain rises on the European nuclear theater. Bulletin of the Atomic Scientists, 39(8):8–10, October 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PP83].

Paine:1983:FC

[Pai83d] Christopher Paine. The freeze and its critics. Bulletin of the Atomic Scientists, 39(4):5–8, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1983:FVT

[Pai83e] Christopher Paine. Freeze verification: time for a fresh approach. Bulletin of the Atomic Scientists, 39(1):6–9, January 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1983:WBB

[Pai83f] Christopher Paine. From Washington: Breakdown on the builddown. Bulletin of the Atomic Scientists, 39(10):4–6, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1983:MTD

[Pai83g] Christopher Paine. MX: too dense for Congress. Bulletin of the Atomic Scientists, 39(2):4–6, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1985:DC

[Pai85a] Christopher Paine. Disarming Congress. Bulletin of the Atomic Scientists, 41(6):6–7, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1985:WRS

[Pai85b] Christopher Paine. Washington report: Senator Nunn’s shell game. Bulletin of the Atomic Scientists, 41(2):5–8, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1985:WRL

[Pai85c] Christopher Paine. Washington report: The low road to arms control. Bulletin of the Atomic Scientists, 41(11):6–7, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1985:WRO

[Pai85d] Christopher Paine. Washington report: The “other nations” speak up. Bulletin of the Atomic Scientists, 41(4):6–7, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1985:LAC

[Pai85e] Christopher E. Paine. Lobbying for arms control. Bulletin of the Atomic Scientists, 41(7):125–130, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hog86].

Paine:1985:WRF

[Pai85f] Christopher E. Paine. Washington report: Fuzzy safeguards for U.S.–China deal. Bulletin of the Atomic Scientists, 41(9):6–7, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:2002:ACM

[Pai02] Christopher Paine. Article commentary: The Moscow Treaty: Making matters worse. Bulletin of the Atomic Scientists, 58 (6):19–21, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/6/19.

Paine:2003:IRP

[Pai03] Christopher Paine. It really is the pits. Bulletin of the Atomic Scientists, 59(5):73, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/5/73.

Paine:2004:BDW

[Pai04a] Christopher Paine. Bush’s “Dirty War”. Bulletin of the Atomic Scientists, 60(6):78–80, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/6/78.

Paine:2004:MH

[Pai04b] Christopher Paine. Down the memory hole. Bulletin of the Atomic Scientists, 60(2):85, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/85.

Paine:2004:UF

[Pai04c] Christopher Paine. A not unreasonable failure? Bulletin of the Atomic Scientists, 60(5):72–73, September/October 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/5/72.

Paine:2004:PP

[Pai04d] Christopher Paine. The party of preemption. Bulletin of the Atomic Scientists, 60(1):75–76, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/1/75.

Paine:2004:WW

[Pai04e] Christopher Paine. Weaponeers of waste. Bulletin of the Atomic Scientists, 60(3):71–72, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/3/71.

Paine:2004:WOM

[Pai04f] Christopher Paine. When one minute = $3.7 billion. Bulletin of the Atomic Scientists, 60(4):75–76, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/75.

Paine:2005:TD

[Pai05] Christopher Paine. Time of delusion. Bulletin of the Atomic Scientists, 61(1):??, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/1/76.

Pajes:1952:LEP

[Paj52] Wolf S. Pajes. Letter to the Editor: Plagues on scientists. Bulletin of the Atomic Scientists, 8(2):65, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Palfrey:1955:ASP

[Pal55] John G. Palfrey. The AEC security program: Past and present. Bulletin of the Atomic Scientists, 11(4):131–133, April 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Palmedo:1964:DFF

[Pal64a] Philip F. Palmedo. The debate on the “Force de Frappe” takes shape. Bulletin of the Atomic Scientists, 20(6):29–30, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Palmer:1964:BBC

[Pal64b] Norman D. Palmer. Books: Quiet Crisis in India: Economic Development and American Policy, by John P. Lewis; The Politics of Scarcity: Public Pressure and Political Response in India, by Myron Weiner. Bulletin of the Atomic Scientists, 20(10):24– 25, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Palme:1972:OE

[Pal72] Olof Palme. The outrage of ecocide. Bulletin of the Atomic Scientists, 28(7):44–45, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Tra72a].

Palladino:1976:LDZ

[Pal76] N. J. Palladino. Letter: Defends zirconium. Bulletin of the Atomic Scientists, 32(3):5, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gul75].

Pal:1977:VVH

[Pal77a] Yash Pal. A visitor to the village: How an American envoy from space demonstrated the great potential of satellite broadcasting. Bulletin of the Atomic Scientists, 33(1):55–56, January 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Palmer:1977:LIW

[Pal77b] L. Palmer. Letter: Inevitability of war. Bulletin of the Atomic Scientists, 33(10):2, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cla77].

Paldy:1984:PSM

[Pal84] Lester G. Paldy. Physical sciences, mathematics and engineering. Bulletin of the Atomic Scientists, 40(10):7s–8s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Palevsky:2000:AFD

[Pal00] Mary Palevsky. Atomic Fragments: a Daughter’s Questions. University of California Press, Berkeley, CA, USA, 2000. ISBN 0-520-22055-2 (hardcover), 0-250-22055-2 (hardcover). xiv + 289 pp. LCCN QC773.3.U5 P35 2000. URL http:

//catdir.loc.gov/catdir/bios/ucal052/99087422.html;
http://catdir.loc.gov/catdir/description/ucal042/99087422.
html.

Pala:2001:DCN

[Pal01] Christopher Pala. Defense conversion: and now for the really big one. Bulletin of the Atomic Scientists, 57(2):20–22, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/2/20.

Palevsky:2002:TTB

[Pal02] Mary Palevsky. Teller on Teller: Book review: Memoirs: A Twentieth-Century Journey in Science and Politics, by Edward Teller and Judith L. Shoolery. Bulletin of the Atomic Scientists, 58(1):64–67, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/1/64;   http://www.
highbeam.com/doc/1G1-81891405.html.

Paltsev:2016:CGR

[Pal16] Sergey Paltsev. The complicated geopolitics of renewable energy. Bulletin of the Atomic Scientists, 72(6):390–395, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pancoast:1954:PFP

[Pan54] Omar B. Pancoast, Jr. The “Point Four” policy. Bulletin of the Atomic Scientists, 10(3):87–92, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Panofsky:1963:BBW

[Pan63] W. K. H. Panofsky. Books: Studies of War, by P. M. S. Blackett. Bulletin of the Atomic Scientists, 19(3):24–25, March 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [AG63, You63a].

Panofsky:1971:RSA

[Pan71] W. K. H. Panofsky. Roots of the strategic arms race: Ambiguity and ignorance. Bulletin of the Atomic Scientists, 27(6):15–20, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Str72].

Panofsky:1981:SRP

[Pan81] W. K. H. Panofsky. Special report from Paris: Science, technology and the arms buildup — I. Bulletin of the Atomic Scientists, 37 (6):48–54, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Panofsky:1984:CPA

[Pan84] Wolfgang K. H. Panofsky. Commentary: Putting the argument to rest. Bulletin of the Atomic Scientists, 40(6):49, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra84a].

Paneth:1986:LRU

[Pan86a] Donald Paneth. Letter: Reforging U.S. journalism. Bulletin of the Atomic Scientists, 42(1):53, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dor85a].

Panofsky:1986:ACN

[Pan86b] Wolfgang K. H. Panofsky. Arms control: necessary process. Bulletin of the Atomic Scientists, 42(3):35–38, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ano86s].

Panofsky:1988:ASL

[Pan88] Wolfgang K. H. Panofsky. After the summit: Limited success, limitless prospects. Bulletin of the Atomic Scientists, 44(2):34– 36, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PNWS:1991:NWS

[Pan91] Panel on Nuclear Weapons Safety. Nuclear weapon safety: How safe is safe? Bulletin of the Atomic Scientists, 47(3):35–40, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Panofsky:1992:DWM

[Pan92] Wolfgang K. H. Panofsky. Destroy weapons; make electricity. Bulletin of the Atomic Scientists, 48(4):15–17, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Panofsky:1998:LVO

[Pan98a] W. K. H. Panofsky. Letter: A viable option. Bulletin of the Atomic Scientists, 54(3):3–4, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pan98b].

Panofsky:1998:R

[Pan98b] Wolfgang K. H. Panofsky. In response: Mixed message. Bulletin of the Atomic Scientists, 54(2):45, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98] and comment [Pan98a].

Panofsky:2000:WKH

[Pan00] W. K. H. Panofsky. W. K. H. Panofsky: We’ll be sorry. Bulletin of the Atomic Scientists, 56(1):47–50, January/February 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/1/47.

Panofsky:2007:PTM

[Pan07a] W. K. H. Panofsky. Peace talk: My life negotiating science and policy. Bulletin of the Atomic Scientists, 63(6):48–53, November/ December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
6/48.

Panofsky:2007:DD

[Pan07b] Wolfgang K. H. Panofsky. A damaging designation. Bulletin of the Atomic Scientists, 63(1):37–39, January/February 2007. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/63/1/37.

Panda:2022:SDC

[Pan22] Ankit Panda. Sure, deter China — but manage risk with North Korea, too. Bulletin of the Atomic Scientists, 78(2):73– 77, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2038884.

Pan:2023:DRN

[Pan23] Yanliang Pan. Diversification from Russian nuclear fuel requires market-oriented solutions. Bulletin of the Atomic Scientists, 80 (1):41–48, 2023. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2023.2293577.

Pape:1988:BRB

[Pap88] Robert A. Pape. Book review: A Thoroughly Efficient Navy, by William M. Kaufmann. Bulletin of the Atomic Scientists, 44 (2):47, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pape:1998:GOW

[Pap98] Robert Pape. Guest opinion: A workable policy on Iraq. Bulletin of the Atomic Scientists, 54(3):6, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parker:1946:PUA

[Par46a] John C. Parker. The public utilities and atomic power. Bulletin of the Atomic Scientists, 1(7):12, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parsons:1946:NSL

[Par46b] Talcott Parsons. National science legislation: Part I: An historical review. Bulletin of the Atomic Scientists, 2(9–10):7–9, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parsons:1947:NSL

[Par47] Talcott Parsons. National science legislation … Part 2: The case for the social sciences. Bulletin of the Atomic Scientists, 3(1):3– 5, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parkman:1953:NMN

[Par53] Henry Parkman. Nonmilitary measures of national defense. Bulletin of the Atomic Scientists, 9(7):259–260, 279, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parkinson:1962:PLM

[Par62] C. Northcote Parkinson. Parkinson’s Laws in medical research. Bulletin of the Atomic Scientists, 18(9):23–24, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PP63].

Park:1966:PHS

[Par66a] Richard Park. Project Harbor summary report. Bulletin of the Atomic Scientists, 22(2):23, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parsegian:1966:FAT

[Par66b] V. L. Parsegian. Forum: Atomic transition — to what? Bulletin of the Atomic Scientists, 22(1):23–26, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Spi66c].

Parsegian:1966:LER

[Par66c] V. L. Parsegian. Letter to the Editor: On the role of government laboratories. Bulletin of the Atomic Scientists, 22(7):35– 36, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spi66b].

Parsegian:1971:NGA

[Par71] V. L. Parsegian. New goals for atomic energy. Bulletin of the Atomic Scientists, 27(8):2–7, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parker:1973:RRE

[Par73] Eugene N. Parker. Research review: Eugene Parker on the solar wind, magnetic fields and earth weather. Bulletin of the Atomic Scientists, 29(5):25–30, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parsegian:1978:LHS

[Par78a] V. L. Parsegian. Letter: Here to stay. Bulletin of the Atomic Scientists, 34(6):3–4, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sau78].

Parsegian:1978:FND

[Par78b] V. Laurence Parsegian. On facing the nuclear dilemma. Bulletin of the Atomic Scientists, 34(2):4–5, February 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Sau78, Fie78].

Parker:1980:BRB

[Par80] Frank L. Parker. Book review: Nuclear Disaster in the Urals, by Zhores S. Medvedev. Bulletin of the Atomic Scientists, 36(4):

49–50, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parker:1983:CLJ

[Par83] Paul M. Parker. Commentary: Lyman J. Briggs. Bulletin of the Atomic Scientists, 39(3):63, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Oli82] and comment [Phi84a].

Park:1985:ILS

[Par85] Robert L. Park. Intimidation leads to self-censorship in science. Bulletin of the Atomic Scientists, 41(3):22–25, March 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Partridge:1986:BRB

[Par86] Ernest Partridge. Book review: Nuclear Weapons and the Future of Humanity, edited by Avner Cohen and Steven Lee. Bulletin of the Atomic Scientists, 42(9):50–52, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parrott:1987:SDM

[Par87] Bruce Parrott. The Soviet debate on missile defense. Bulletin of the Atomic Scientists, 43(3):9–12, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parnas:1988:LSS

[Par88a] David L. Parnas. Letter: SDI sidebar errors. Bulletin of the Atomic Scientists, 44(5):53, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gur88b].

Parsons:1988:BRB

[Par88b] Howard L. Parsons. Book review: The Nuclear Predicament: Explorations in Soviets Ideology, by Stephen Shenfield. Bulletin of the Atomic Scientists, 44(4):51–52, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parthemore:2017:ACW

[Par17] Christine Parthemore. The ambiguity challenge: Why the world needs a multilateral nuclear cruise missile agreement. Bulletin of the Atomic Scientists, 73(3):154–158, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1315919.

Paschkis:1951:LEM

[Pas51] Victor Paschkis. Letter to the Editor: The mobilization of scientists. Bulletin of the Atomic Scientists, 7(4):108, April 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano50z].

Paschkis:1955:SRP

[Pas55a] Victor Paschkis. The scientist’s responsibility: A pacifist view. Bulletin of the Atomic Scientists, 11(7):265–266, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Passin:1955:JHB

[Pas55b] Herbert Passin. Japan and the H-bomb. Bulletin of the Atomic Scientists, 11(8):289–292, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paschkis:1966:LES

[Pas66] Victor Paschkis. Letter to the Editor: “Science”. Bulletin of the Atomic Scientists, 22(4):30–31, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [dSP65].

Paschkis:1974:CST

[Pas74] Victor Paschkis. Communications: Science and technology. Bulletin of the Atomic Scientists, 30(3):2–3, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor73, Smi74d].

Paskusz:1976:LBW

[Pas76] G. F. Paskusz. Letter: Black and white. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Paschkis:1977:LPM

[Pas77] Victor Paschkis. Letter: Public must decide. Bulletin of the Atomic Scientists, 33(6):8, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kan77a, AB77].

Patel:1962:ECD

[Pat62] Surendra J. Patel. Economic consequences of disarmament. Bulletin of the Atomic Scientists, 18(9):7–12, November 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paterson:1975:MPL

[Pat75] Arthur E. Paterson III. Marine pollution and the Law of the Sea. Bulletin of the Atomic Scientists, 31(10):48–50, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1978:CLE

[Pat78a] Walter C. Patterson. Commentary: From London: Energy conservation: ‘not doing without but doing more’. Bulletin of the Atomic Scientists, 34(10):6–7, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1978:CLP

[Pat78b] Walter C. Patterson. Commentary: From London: “Put it somewhere else!”. Bulletin of the Atomic Scientists, 34(9):6–7, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1978:WRN

[Pat78c] Walter C. Patterson. The Windscale Report: a nuclear apologia. Bulletin of the Atomic Scientists, 34(6):44–49, June 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patin:1979:BRB

[Pat79a] Henry Patin. Book review: Détente or Debacle, edited by Fred Warner Neal. Bulletin of the Atomic Scientists, 35(7):46, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1979:CAA

[Pat79b] Walter C. Patterson. Commentary from Austria: Austria’s nuclear referendum. Bulletin of the Atomic Scientists, 35(1):6–7, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [NW79].

Patterson:1979:CLC

[Pat79c] Walter C. Patterson. Commentary from London: Conservation cornucopia. Bulletin of the Atomic Scientists, 35(5):42–44, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1979:CLH

[Pat79d] Walter C. Patterson. Commentary from London: Harrisburg is überall. Bulletin of the Atomic Scientists, 35(6):9–11, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1979:CLN

[Pat79e] Walter C. Patterson. Commentary from London: Nuclear polarization. Bulletin of the Atomic Scientists, 35(9):6–7, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1979:CLP

[Pat79f] Walter C. Patterson. Commentary from London: Positive thinking. Bulletin of the Atomic Scientists, 35(10):6–7, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1979:GH

[Pat79g] Walter C. Patterson. Gorleben hearings. Bulletin of the Atomic Scientists, 35(6):11, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1980:CLB

[Pat80a] Walter C. Patterson. Commentary from London: Breeders interbreeding. Bulletin of the Atomic Scientists, 36(9):5–7, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1980:CLD

[Pat80b] Walter C. Patterson. Commentary from London: Don’t just do something — stand there! Bulletin of the Atomic Scientists, 36(2):7–8, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1980:CLF

[Pat80c] Walter C. Patterson. Commentary from London: Future imperfect. Bulletin of the Atomic Scientists, 36(4):11–13, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1980:CLN

[Pat80d] Walter C. Patterson. Commentary from London: The nuclear civil war. Bulletin of the Atomic Scientists, 36(5):13–14, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1980:CLO

[Pat80e] Walter C. Patterson. Commentary from London: The Open University tackles control of technology. Bulletin of the Atomic Scientists, 36(3):56–57, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1981:BRB

[Pat81a] Walter C. Patterson. Book reviews: Energy in France: Planning, Politics and Policy, by N. J. D. Lucas; Decision-making for Energy Futures by David Pearce, Lynn Edwards and Geoff Beuret; Renewable Energy, by Bent Sørensen; Solar versus Nuclear: Choosing Energy Futures. Bulletin of the Atomic Scientists, 37(3):46–47, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1981:CLN

[Pat81b] Walter C. Patterson. Commentary from London: Nuclear street fighting. Bulletin of the Atomic Scientists, 37(1):59–61, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1981:CLT

[Pat81c] Walter C. Patterson. Commentary from London: Twenty-five years in a quandary. Bulletin of the Atomic Scientists, 37(8):56– 58, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1982:BRB

[Pat82a] Walter C. Patterson. Book review: Future Energy Consumption of the Third World, by Markus Fritz. Bulletin of the Atomic Scientists, 38(2):37–38, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1982:PS

[Pat82b] Walter C. Patterson. Plutonium for sale. Bulletin of the Atomic Scientists, 38(5):55–56, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1983:BRB

[Pat83] Walter C. Patterson. Book review: The Cult of the Atom, by Daniel Ford. Bulletin of the Atomic Scientists, 39(7):45–46, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1984:RS

[Pat84] Walter C. Patterson. A report on Sizewell. Bulletin of the Atomic Scientists, 40(6):43–44, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1986:BRP

[Pat86a] Walter C. Patterson. Breeder reactor politics in Europe. Bulletin of the Atomic Scientists, 42(5):37–40, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1986:COS

[Pat86b] Walter C. Patterson. Chernobyl — the official story. Bulletin of the Atomic Scientists, 42(9):34–36, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1986:CES

[Pat86c] Walter C. Patterson. Chernobyl: the emerging story: Chernobyl: worst but not first. Bulletin of the Atomic Scientists, 42(7):43– 45, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patterson:1987:LOC

[Pat87] E. F. Patterson. Letter: Origins of the Cold War. Bulletin of the Atomic Scientists, 43(6):62, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Boy87] and comment [Nin88].

Patterson:1988:JPP

[Pat88] Walter C. Patterson. Japan’s perilous plutonium flights. Bulletin of the Atomic Scientists, 44(4):9–11, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Per88a, TN90, BP90].

Patterson:1990:RGB

[Pat90a] Walter C. Patterson. Report: Great Britain: Thatcher’s failed romance with nuclear power. Bulletin of the Atomic Scientists, 46 (3):10–12, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Patton:1990:LFS

[Pat90b] Garry A. Patton. Letter: FEMA should stick to natural disasters, says Santa Cruz. Bulletin of the Atomic Scientists, 46(3):53, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack89h].

Pauling:1952:DR

[Pau52a] Linus Pauling. The decision reversed. Bulletin of the Atomic Scientists, 8(7):256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Pauling:1952:MEO

[Pau52b] Linus Pauling. My efforts to obtain a passport. Bulletin of the Atomic Scientists, 8(7):253–255, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Pauling:1957:AAS

[Pau57] Linus Pauling. An appeal by American scientists to the governments and people of the world. Bulletin of the Atomic Scientists, 13(7):264–266, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The appeal is signed by 2173 scientists, plus many nonscientists.

Pauling:1962:GEW

[Pau62] Linus Pauling. Genetic effects of weapons tests. Bulletin of the Atomic Scientists, 18(10):15–18, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pauling:1963:FGE

[Pau63] Linus Pauling. Forum: Genetic effects. Bulletin of the Atomic Scientists, 19(7):30–32, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RGR63].

Pauling:1967:PEP

[Pau67] Linus Pauling. Peace on Earth: The position of the scientists. Bulletin of the Atomic Scientists, 23(8):46–48, October 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pauling:1970:TGP

[Pau70] Linus Pauling. Tamplin–Gofman, Pauling and the AEC: Genetic and somatic effects of high-energy radiation. Bulletin of the Atomic Scientists, 26(7):3–5, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TG70].

Pauling:1977:NFA

[Pau77] Linus Pauling. Nuclear fission is not the answer. Bulletin of the Atomic Scientists, 33(3):66, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pauling:1979:PNP

[Pau79] Linus Pauling. Perspective: Nobel Peace Laureates: 1959. Bulletin of the Atomic Scientists, 35(6):1, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 10 December 1963, Oslo, Norway.

Paul:1988:LHI

[Pau88] Liz Paul. Letter: Halt Idaho’s SIS. Bulletin of the Atomic Scientists, 44(3):56, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack88n].

Paulikas:2005:LLP

[Pau05] Steven Paulikas. Lithuania: Losing power. Bulletin of the Atomic Scientists, 61(2):12–14, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/2/12.

Pavlov:1998:NBV

[Pav98a] Stefan Pavlov. NATO: Bulgaria in a vise. Bulletin of the Atomic Scientists, 54(1):28–31, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pavlov:1998:SC

[Pav98b] Stefan Pavlov. Sofia’s choice. Bulletin of the Atomic Scientists, 54(3):52–57, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Payne:1970:CUN

[Pay70] Buryl Payne. Communications: United Nations tax. Bulletin of the Atomic Scientists, 26(9):40, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Payette:1984:CNG

[Pay84] Thomas J. Payette. Commentary: Nerve gas. Bulletin of the Atomic Scientists, 40(5):62, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa83a].

Parrish:1968:FWS

[PB68] John B. Parrish and Jean S. Block. The future of women in science and engineering. Bulletin of the Atomic Scientists, 24(5): 46–49, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Palmer:1984:LIS

[PB84] George Palmer and Dan I. Bolef. Laser isotope separation: the plutonium connection. Bulletin of the Atomic Scientists, 40(3): 26–31, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [IPB84, WPB84].

Phillips:1992:LPN

[PB92] Alan F. Phillips and Joanna Santa Barbara. Letter: Preemptive nonproliferation. Bulletin of the Atomic Scientists, 48(5):47, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AH92e].

Podvig:2001:RSN

[PB^{+}01] P. L. (Pavel Leonardovich) Podvig, Oleg Bukharin, et al., editors. Russian strategic nuclear forces. MIT Press, Cambridge, MA, USA, 2001. ISBN 0-262-16202-4 (hardcover). xxi + 692 pp. LCCN UA776.R37 R87 2001.

Phelps:1955:TPS

[PBG55] John B. Phelps, Ralph S. Brown, Jr., and S. A. Goudsmit. Toward a positive security program. Bulletin of the Atomic Scientists, 11(4):165–169, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Paine:1992:HSW

[PC92a] Christopher Paine and Thomas B. Cochran. A how-to session on warhead destruction. Bulletin of the Atomic Scientists, 48(1): 14, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Paine:1992:LTM

[PC92b] Christopher Paine and Thomas B. Cochran. So little time, so many weapons, so much to do. Bulletin of the Atomic Scientists, 48(1):12–16, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1949:FS

[PD49] Rudolf Peierls and Sir Henry Dale. Freedom of science. Bulletin of the Atomic Scientists, 5(4):106–109, April 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkens:1989:CNA

[PD89] Roy Perkens, Jr. and Sidney Drell. Counterpoint: Nuclear abolition: would cheaters count? Bulletin of the Atomic Scientists, 45 (10):16–17, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pidgeon:2012:NRE

[PD12] Nick Pidgeon and Christina C. Demski. From nuclear to renewable: Energy system transformation and public attitudes. Bulletin of the Atomic Scientists, 68(4):41–51, July/August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/4/41.

Price:1989:LTB

[PE89] H. Edward Price, Jr. and Jack F. Evernden. Letters: Test ban, round three. Bulletin of the Atomic Scientists, 45(1):64, 66, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eve88a].

Park:2024:EIU

[PE24] Sulgiye Park and Rodney C. Ewing. Environmental impacts of underground nuclear weapons testing. Bulletin of the Atomic Scientists, 80(2):102–111, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2314439.

Pearson:1958:PP

[Pea58] Lester Pearson. Peace and prosperity. Bulletin of the Atomic Scientists, 14(6):214, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peattie:1984:NU

[Pea84] Lisa Peattie. Normalizing the unthinkable. Bulletin of the Atomic Scientists, 40(3):32–36, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bre84].

Pearson:1985:OUN

[Pea85] Ruth Pearson. 1985 outlook: United Nations turns 40. Bulletin of the Atomic Scientists, 41(1):18–19, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pearson:1988:UCU

[Pea88] Ruth Pearson. U.N. cries “uncle”. Bulletin of the Atomic Scientists, 44(8):36–39, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pearson:1990:LBT

[Pea90] J. M. Pearson. Letter: Better than PALs for submarines. Bulletin of the Atomic Scientists, 46(2):45, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zim89b].

Pearson:1991:RUN

[Pea91] Ruth Pearson. Report: United Nations: Beyond the ‘chartered’ path. Bulletin of the Atomic Scientists, 47(10):8–10, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pearson:1992:CC

[Pea92a] Ruth Pearson. Castro’s charisma. Bulletin of the Atomic Scientists, 48(10):48, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pearson:1992:CDJ

[Pea92b] Ruth Pearson. Cuba’s double jeopardy. Bulletin of the Atomic Scientists, 48(10):44–50, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pearson:1992:TR

[Pea92c] Ruth Pearson. Two revolutions. Bulletin of the Atomic Scientists, 48(10):49, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pearson:1999:UNA

[Pea99] Ruth Pearson. United Nations: Angst on the East River. Bulletin of the Atomic Scientists, 55(6):??, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/6/19.

Pecora:1970:SQO

[Pec70] W. T. Pecora. Science and the quality of our environment. Bulletin of the Atomic Scientists, 26(8):20–23, October 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rum71a].

Peel:1978:BBC

[Pee78] J. D. Y. Peel. Books: Culture and its Creators: Essays in honor of Edward Shils, by Joseph Ben-David and Terry Nichols Clark. Bulletin of the Atomic Scientists, 34(3):61–62, March 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1950:BAS

[Pei50] Rudolf E. Peierls. The British Atomic Scientists’ Association. Bulletin of the Atomic Scientists, 6(2):59, February 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1952:SBE

[Pei52] R. E. Peierls. Some British experiences. Bulletin of the Atomic Scientists, 8(7):229–230, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Peierls:1953:GEB

[Pei53] R. E. Peierls. Guest editorial: Basic science and the Cold War. Bulletin of the Atomic Scientists, 9(3):66–67, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1969:SPA

[Pei69] Sir Rudolf E. Peierls. The scientist in public affairs: Between the ivory tower and the arena. Bulletin of the Atomic Scientists, 25 (9):28–30, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1970:BAA

[Pei70] Sir Rudolf E. Peierls. Britain in the Atomic Age. Bulletin of the Atomic Scientists, 26(6):40–46, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1980:BRE

[Pei80] Sir Rudolf Peierls. Book review: Energy from Heaven and Earth, by Edward Teller. Bulletin of the Atomic Scientists, 36(5):53–54, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1982:LNW

[Pei82] Sir Rudolf Peierls. Limited nuclear war? Bulletin of the Atomic Scientists, 38(5):2, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1984:BND

[Pei84] Sir Rudolf Peierls. Breaking the nuclear deadlock: First strike fiction. Bulletin of the Atomic Scientists, 40(1):50, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1985:RBP

[Pei85] Rudolf Peierls. Reflections of a British participant. Bulletin of the Atomic Scientists, 41(7):27–29, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pell:1964:PPG

[Pel64] Claiborne Pell. Pell’s proposals on Germany. Bulletin of the Atomic Scientists, 20(2):9, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Penney:1952:MBE

[Pen52] Sir William Penney. The Monte Bello explosion. Bulletin of the Atomic Scientists, 8(9):295–296, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Penkrot:1977:LAP

[Pen77a] Jack Penkrot. Letter: Another plutonium source? Bulletin of the Atomic Scientists, 33(7):3–4, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste77a].

Penkrot:1977:LCF

[Pen77b] Jack Penkrot. Letter: Cowering in fear. Bulletin of the Atomic Scientists, 33(1):6, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Flo76a] and comment [Sut77].

Penberthy:1984:CWP

[Pen84] Larry Penberthy. Commentary: Who pays for radioactive rubbish? Bulletin of the Atomic Scientists, 40(9):57, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Abb84b].

Pentz:1998:LOC

[Pen98] Linda Pentz. Letter: Not one cent, red or otherwise. Bulletin of the Atomic Scientists, 54(6):3, 64, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [S^{+}98].

Peppard:2015:PFF

[Pep15] Christiana Z. Peppard. Pope Francis and the fourth era of the Catholic Church’s engagement with science. Bulletin of the Atomic Scientists, 71(5):31–39, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/71/5/31.full.pdf+html.

Perrin:1955:AEF

[Per55] François Perrin. Atomic energy in France. Bulletin of the Atomic Scientists, 11(3):92–93, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perl:1971:SI

[Per71] Martin L. Perl. SALT and its illusions. Bulletin of the Atomic Scientists, 27(10):7–12, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkins:1978:ANE

[Per78] Dwight H. Perkins. Asia’s new economic development. Bulletin of the Atomic Scientists, 34(8):11–18, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkins:1985:DI

[Per85] Raymond K. Perkins, Jr. Deterrence is immoral. Bulletin of the Atomic Scientists, 41(2):32–34, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [MS85].

Perlo:1987:LEM

[Per87a] Victor Perlo. Letter: Excluding the most excludable. Bulletin of the Atomic Scientists, 43(6):59–60, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kal87].

Perrow:1987:BRB

[Per87b] Charles Perrow. Book review: Averting Catastrophe: Strategies for Regulating Risky Technologies, by Joseph G. Morone and Edward J. Woodhouse. Bulletin of the Atomic Scientists, 43(3): 48–50, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkins:1988:LES

[Per88a] James Perkins. Letter: Elegant solution. Bulletin of the Atomic Scientists, 44(6):54, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pat88].

Perkovich:1988:WC

[Per88b] George Perkovich. So who’s counting? Bulletin of the Atomic Scientists, 44(7):17, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pery:1990:BHH

[Per90] Todd Pery. Bulletin: Hands off health studies. Bulletin of the Atomic Scientists, 46(4):4–5, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkovich:1991:BRP

[Per91a] George Perkovich. Book review: Psychological warfare: The Imaginary War, by Mary Kaldor. Bulletin of the Atomic Scientists, 47(8):41–43, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkovich:1991:NWO

[Per91b] George Perkovich. New world order: Put nuclear weapons on the agenda. Bulletin of the Atomic Scientists, 47(5):21–23, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkovich:1992:PWC

[Per92a] George Perkovich. Perspective: Weapons complexes v. democracy. Bulletin of the Atomic Scientists, 48(5):16–17, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perry:1992:LEB

[Per92b] Edward C. Perry. Letter: Earlier “blueprint”. Bulletin of the Atomic Scientists, 48(3):47, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RD91].

Perkovich:1994:BRB

[Per94] George Perkovich. Book review: A Preponderance of Power, by Melvyn Leffler. Bulletin of the Atomic Scientists, 50(3):58, 61, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkovich:1995:BRS

[Per95] George Perkovich. Book review: South Asian instability: Indo-Pak Nuclear Standoff: The Role of the United States, by P. R. Chari. Bulletin of the Atomic Scientists, 51(5):53–56, September/ October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perlstein:1997:AQ

[Per97] Rick Perlstein. Author’s query. Bulletin of the Atomic Scientists, 53(3):59, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perkovich:2000:BQB

[Per00] George Perkovich. Bhabha’s quest for The bomb. Bulletin of the Atomic Scientists, 56(3):54–63, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/3/54.

Perkovich:2002:IA

[Per02] George Perkovich. India’s ambitions. Bulletin of the Atomic Scientists, 58(2):66–68, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/2/66.

Perrow:2011:FIA

[Per11] Charles Perrow. Fukushima and the inevitability of accidents. Bulletin of the Atomic Scientists, 67(6):44–52, November/ December 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
6/44.

Perrow:2013:NDH

[Per13] Charles Perrow. Nuclear denial: From Hiroshima to Fukushima. Bulletin of the Atomic Scientists, 69(5):56–67, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/69/5/56.

Perry:2020:HUD

[Per20] William J. Perry. How a US Defense Secretary came to support the abolition of nuclear weapons. Bulletin of the Atomic Scientists, 76(6):290–293, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1846408.

Pessen:1987:LBP

[Pes87] Edward Pessen. Letter: Baruch’s ploy. Bulletin of the Atomic Scientists, 43(4):55, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fal87].

Peter:1950:LLW

[Pet50] Marc Peter, Jr. Lessons from the last war. Bulletin of the Atomic Scientists, 6(8–9):252–255, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Peterson:1954:ME

[Pet54] Val Peterson. Mass evacuation. Bulletin of the Atomic Scientists, 10(7):294–295, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Petrik:1961:RCC

[Pet61] Ester Petrik. Reflections and comments: Controversy on Kahn. Bulletin of the Atomic Scientists, 17(7):293–294, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar61d, Kea61, Mar61c].

Petersen:1963:PNB

[Pet63] Aage Petersen. The philosophy of Niels Bohr. Bulletin of the Atomic Scientists, 19(7):8–14, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peters:1978:LBR

[Pet78] James Peters. Letter: A big request. Bulletin of the Atomic Scientists, 34(5):6, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Petrod:1981:CCF

[Pet81] Martine Petrod. Commentary: Copenhagen Foundation Against Nuclear Tests. Bulletin of the Atomic Scientists, 37(7):64, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peterson:1992:LWM

[Pet92] David Peterson. Letter: We have met the enemy and …. Bulletin of the Atomic Scientists, 48(5):47, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo91b].

Petrarolha:1996:BMW

[Pet96] Fábio L. S. Petrarolha. Brazil: the meek want the earth now. Bulletin of the Atomic Scientists, 52(6):20–29, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Petryna:2011:CSP

[Pet11] Adriana Petryna. Chernobyl’s survivors: Paralyzed by fatalism or overlooked by science? Bulletin of the Atomic Scientists, 67 (2):30–37, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/2/30.

Porter:1975:BBA

[PF75] Alan L. Porter and Steven H. Flajser. Books: Society and the Assessment of Technology, by François Hetman; Environmental Harassment OR Technology Assessment, by Derek Medford; Technology Assessment in a Dynamic Environment, edited by M. J. Cetron and B. Bartocha; Technology Assessment and Quality of Life, edited by G. J. Stöber and D. Schumacher. Bulletin of the Atomic Scientists, 31(5):47–48, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Perlo:1982:CRC

[PF82] Victor Perlo and Bernard T. Feld. Commentary: Rereading the Clock issue. Bulletin of the Atomic Scientists, 38(8):58–59, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel82k].

Pfeil:1951:MDH

[Pfe51] P. C. L. Pfeil. Metallurgy [Division at Harwell]. Bulletin of the Atomic Scientists, 7(7–8):231–232, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pfeufer:1952:LEB

[Pfe52] Philip Paul Pfeufer. Letter to the Editor: Science and the Catholic Church. Bulletin of the Atomic Scientists, 8(6):209, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pfeiffer:1982:ORH

[Pfe82] E. W. Pfeiffer. Operation Ranch Hand: the U.S. herbicide program. Bulletin of the Atomic Scientists, 38(5):20–24, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parthemore:2018:GRP

[PFW18] Christine Parthemore, Francesco Femia, and Caitlin Werrell. The global responsibility to prepare for intersecting climate and nuclear risks. Bulletin of the Atomic Scientists, 74(6):374–378, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1533186.

Petitpierre:1979:CSI

[PG79] G. Petitpierre and B. Giovannini. Commentary from Switzerland: 1979 initiative. Bulletin of the Atomic Scientists, 35(5):49, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prudori:1984:CQV

[PG84] Tony Prudori and Robert Gomer. Commentary: Quo vadis? Bulletin of the Atomic Scientists, 40(6):55–56, June/July 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom84b] and comment [Hen84].

Plotkin:1991:LTT

[PG91] Sheldon C. Plotkin and Daniel Gibson. Letters: Transmutation — tried but not true. Bulletin of the Atomic Scientists, 47(8): 47, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gib91a].

Poschmann:1991:LFI

[PGL91] Finn Poschmann, George G. Giddings, and Donald B. Louria. Letter: Food irradiation: readers zap back. Bulletin of the Atomic Scientists, 47(2):46–47, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lou90].

Phelps:1956:BBS

[Phe56] John B. Phelps. Books: Case Studies in Personal Security, collected under the direction of Adam Yarmolinsky. Bulletin of the Atomic Scientists, 12(1):25, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Phelps:1961:DAW

[Phe61] John B. Phelps. The danger of accidental war. Bulletin of the Atomic Scientists, 17(4):146–148, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Phelps:1963:BBS

[Phe63] John Phelps. Books: Fail-Safe, by Eugene Burdick and Harvey Wheeler. Bulletin of the Atomic Scientists, 19(6):28, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Phillips:1954:ARW

[Phi54] Thomas R. Phillips. The atomic revolution in warfare. Bulletin of the Atomic Scientists, 10(8):315–317, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Phillips:1960:UI

[Phi60] Thomas R. Phillips. The U-2 incident. Bulletin of the Atomic Scientists, 16(6):222, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Philip:1984:LJB

[Phi84a] J. R. Philip. On ‘Lyman J. Briggs Revisited’. Bulletin of the Atomic Scientists, 40(3):64, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Par83].

Phillips:1984:CMC

[Phi84b] John N. Phillips. Commentary: Move the clock. Bulletin of the Atomic Scientists, 40(10):35, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lyt85].

Phillips:1993:LMD

[Phi93] Alan F. Phillips. Letter: Moral distinctions. Bulletin of the Atomic Scientists, 49(2):45, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo92l].

Phillips:1994:LMR

[Phi94] Alan F. Phillips. Letter: Meager rationality. Bulletin of the Atomic Scientists, 50(3):60, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo93b].

PSR:1980:OLP

[Phy80] Physicians for Social Responsibility. Open letter to President Jimmy Carter and Chairman Leonid Brezhnev: Danger of nuclear war. Bulletin of the Atomic Scientists, 36(4):16, April 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Plesch:1992:RWS

[PI92] Daniel Plesch and Sandra Ionno. Report: Weapon safety: Bouncing bombs across Britain. Bulletin of the Atomic Scientists, 48 (10):8–9, 52, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pickavance:1951:RBA

[Pic51] T. G. Pickavance. Report on the British Atomic Energy Research Establishment. Bulletin of the Atomic Scientists, 7(7–8):229–230, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pickering:2010:IQH

[Pic10] Thomas R. Pickering. The Iranian quagmire: How to move forward: Position: Confident diplomacy. Bulletin of the Atomic Scientists, 66(6):88–94, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/66/6/88.

Piel:1954:SNF

[Pie54] Gerard Piel. Science and the next fifty years. Bulletin of the Atomic Scientists, 10(1):17–20, 25, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1955:PE

[Pie55] Gerard Piel. The planet Earth. Bulletin of the Atomic Scientists, 11(7):238–243, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1958:SDP

[Pie58] Gerard Piel. Science, disarmament, and peace. Bulletin of the Atomic Scientists, 14(6):217–219, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1959:RML

[Pie59] Gerard Piel. The revolution in man’s labor. Bulletin of the Atomic Scientists, 15(7):278–283, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1960:ED

[Pie60] Gerard Piel. The economics of disarmament. Bulletin of the Atomic Scientists, 16(4):117–122, 126, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fla60].

Pierce:1961:HCS

[Pie61] J. R. Pierce. Hazards of communications satellites. Bulletin of the Atomic Scientists, 17(5–6):181–185, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1962:ICD

[Pie62] Gerard Piel. The illusion of civil defense. Bulletin of the Atomic Scientists, 18(2):2–8, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1963:AA

[Pie63] Gerard Piel. The advent of abundance. Bulletin of the Atomic Scientists, 19(6):2–6, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1966:FFS

[Pie66] Gerard Piel. Federal funds and science education. Bulletin of the Atomic Scientists, 22(5):10–15, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Piel:1968:WFW

[Pie68] Gerard Piel. A world free of want? Bulletin of the Atomic Scientists, 24(1):16–22, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pierre:1973:CNP

[Pie73] Andrew J. Pierre. Communications: Nuclear politics in Britain. Bulletin of the Atomic Scientists, 29(10):3, December 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar73a].

Pietsch:1980:CSA

[Pie80] Stephen Pietsch. Commentary: SALT and the arms race. Bulletin of the Atomic Scientists, 36(3):63, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Div79].

Pierrehumbert:2016:HDL

[Pie16] Raymond Pierrehumbert. How to decarbonize? Look to Sweden. Bulletin of the Atomic Scientists, 72(2):105–111, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pierrehumbert:2019:TNP

[Pie19] Raymond Pierrehumbert. There is no Plan B for dealing with the climate crisis. Bulletin of the Atomic Scientists, 75(5):215– 221, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1654255.

Pierrehumbert:2022:PPB

[Pie22] Raymond Pierrehumbert. Plant power: Burning biomass instead of coal can help fight climate change — but only if done right. Bulletin of the Atomic Scientists, 78(3):125–127, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2062931.

Pifer:2018:WMN

[Pif18] Steven Pifer. Washington–Moscow nuclear verification: Tensions and solutions. Bulletin of the Atomic Scientists, 74(5): 297–304, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1507775.

Pifer:2024:LUR

[Pif24] Steven Pifer. The logic for US ratification of the Comprehensive Nuclear Test Ban Treaty. Bulletin of the Atomic Scientists, 80(2):81–86, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2314433.

Pike:1948:CS

[Pik48] Sumner T. Pike. A Commissioner speaks. Bulletin of the Atomic Scientists, 4(1):15–17, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pike:1996:GOB

[Pik96] John Pike. Guest opinion: Buying votes with B-2s. Bulletin of the Atomic Scientists, 52(3):4, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pikayev:2006:TR

[Pik06] Alexander A. Pikayev. Targeting Russia? Bulletin of the Atomic Scientists, 62(1):63–64, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/63.

Pikayev:2007:UA

[Pik07] Alexander A. Pikayev. Unfair advantage. Bulletin of the Atomic Scientists, 63(4):36, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/4/36.

Pilisuk:1982:GSP

[Pil82] Mark Pilisuk. Games strategists play. Bulletin of the Atomic Scientists, 38(9):13–17, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pinke:1957:RCL

[Pin57] A. S. Pinke. Reader comment: Limitation of fissionable material in weapons. Bulletin of the Atomic Scientists, 13(5):177–178, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pincus:2024:HNW

[Pin24] Walter Pincus. The horrors of nuclear weapons testing. Bulletin of the Atomic Scientists, 80(2):112–117, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2314436.

Piper:1955:LEA

[Pip55] Henry Dan Piper. Letter to the Editor: Are American scientists ashamed of their history? Bulletin of the Atomic Scientists, 11 (2):73, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pistolese:1964:LEM

[Pis64] Francesco Pistolese. Letter to the Editor: More on Lilienthal. Bulletin of the Atomic Scientists, 20(5):29, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lil63].

Pitzer:1952:LE

[Pit52a] Kenneth Pitzer. Letter to the Editor: Power progress too slow. Bulletin of the Atomic Scientists, 8(9):312, December 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pit52b] for the abbreviated speech; the full speech is published in Nucleonics, April 1952.

Pitzer:1952:PPT

[Pit52b] Kenneth S. Pitzer. Power progress too slow. Bulletin of the Atomic Scientists, 8(8):287–288, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [FDC52, Pit52a].

Pittendrigh:1961:BSE

[Pit61] Colin S. Pittendrigh. Biology and space environment. Bulletin of the Atomic Scientists, 17(5–6):206–209, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pittman:1964:CDB

[Pit64] Steuart L. Pittman. Civil defense in a balanced national security. Bulletin of the Atomic Scientists, 20(6):24–26, June 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pittendrigh:1967:BSS

[Pit67] Colin S. Pittendrigh. The biologist in the solar system. Bulletin of the Atomic Scientists, 23(3):4–10, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pitman:1991:BRT

[Pit91] George R. Pitman. Book review: Taking aim: Inventing Accuracy: A Historical Sociology of Nuclear Missile Guidance, by Donald MacKenzie. Bulletin of the Atomic Scientists, 47(8):43– 44, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mil91].

Phelps:1963:TCP

[PK63] John Phelps and Marvin Kalkstein. The N th country problem. Bulletin of the Atomic Scientists, 19(10):50–51, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1953:IDC

[Pla53] William J. Platt. Industrial defense: A community approach. Bulletin of the Atomic Scientists, 9(7):261–264, September 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1957:BBH

[Pla57a] John Rader Platt. Books: Science and Human Life, by J. A. V. Butler. Bulletin of the Atomic Scientists, 13(7):272, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1957:BBN

[Pla57b] John Rader Platt. Books: The Next Hundred Years, by Harrison Brown, James Bonner, and John Weir; Science and Economic Development, by Richard L. Meier; Technology and Social Change, by Francis R. Allen, Hornell Hart, Delbert C. Miller, William F. Ogburn, and Meyer F. Nimkoff. Bulletin of the Atomic Scientists, 13(9):341–342, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1959:BSC

[Pla59a] John Rader Platt. The Search, by C. P. Snow; The Scientists, by Eleazar Lipsky. Bulletin of the Atomic Scientists, 15(6):268, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1959:CC

[Pla59b] John Rader Platt. Competition in creation. Bulletin of the Atomic Scientists, 15(2):82–85, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1960:BBR

[Pla60] John Rader Platt. Books: Basic Research in the Navy. Bulletin of the Atomic Scientists, 16(6):220–221, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1961:SCR

[Pla61] John R. Platt. Social chain reactions. Bulletin of the Atomic Scientists, 17(9):365–369, 386, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1964:BBC

[Pla64a] John R. Platt. Books: Continuities in Cultural Evolution, by Margaret Mead. Bulletin of the Atomic Scientists, 20(10):25– 26, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1964:RDS

[Pla64b] John R. Platt. Research and development for social problems. Bulletin of the Atomic Scientists, 20(6):27–29, June 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1965:NVN

[Pla65] John R. Platt. New views of the nature of man: The Monday lectures. Bulletin of the Atomic Scientists, 21(7):12, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1969:BBB

[Pla69] Joseph B. Platt. Books: Education at the Barricades, by Charles Frankel. Bulletin of the Atomic Scientists, 25(5):53, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1970:HG

[Pla70] John Platt. Hierarchical growth. Bulletin of the Atomic Scientists, 26(9):2–4, 46–48, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1973:VST

[Pla73] Joseph B. Platt. The value of science and technology to human welfare. Bulletin of the Atomic Scientists, 29(8):16–21, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi74c].

Platt:1978:COE

[Pla78] John Platt. Commentary: Our electronic world. Bulletin of the Atomic Scientists, 34(10):49–50, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Platt:1984:CMS

[Pla84] Stanley K. Platt. Commentary: Mutual security. Bulletin of the Atomic Scientists, 40(3):63, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Plesch:1991:RNS

[Ple91] Daniel Plesch. Report: NATO: Same tune, fewer violins. Bulletin of the Atomic Scientists, 47(9):9–10, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Plesch:1996:GOL

[Ple96] Daniel Plesch. Guest opinion: Letting defeatism defeat disarmament. Bulletin of the Atomic Scientists, 52(4):4, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Plous:1986:PNF

[Plo86] Scott Plous. Perspective: No first use: having it both ways. Bulletin of the Atomic Scientists, 42(1):10–11, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [EP86].

Plous:1989:BWB

[Plo89] Scott Plous. Bulletin: When broken arrows show. Bulletin of the Atomic Scientists, 45(10):3–4, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Parsons:1950:BRB

[PMRP50] W. S. Parsons, Hans J. Morgenthau, Robert Redfield, and David Phillips. Book reviews: Modern Arms and Free Men, by Vannevar Bush, Ape and Essence, by Aldous Huxley, This I Do Believe: An American Credo, by David E. Lilienthal. Bulletin of the Atomic Scientists, 6(3):91–96, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Potter:2000:TNL

[PMS00] William C. Potter, Djuro Miljanic, and Ivo Slaus. Tito’s nuclear legacy. Bulletin of the Atomic Scientists, 56(2):63–70, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/2/63.

Polmar:2009:UNA

[PN09] Norman Polmar and Robert S. (Robert Stan) Norris. The U.S. nuclear arsenal: a history of weapons and delivery systems since 1945. Naval Institute Press, Annapolis, MD, USA, 2009. ISBN 1-55750-681-7. xvi + 274 pp. LCCN U264.3 .P65 2009.

Phillips:2016:TAS

[PO16] Fred Young Phillips and Deog-Seong Oh. Technology assessment and the social and human impact of innovation. Bulletin of the Atomic Scientists, 72(6):402–411, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pocha:2005:MUF

[Poc05] Jehangir S. Pocha. Mongolia: Unlikely facilitator. Bulletin of the Atomic Scientists, 61(6):18–20, November/December 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/6/18.

Podvig:2001:MDR

[Pod01] Pavel Podvig. Missile defense: For Russia, little loss, little gain. Bulletin of the Atomic Scientists, 57(6):23–25, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/6/23.

Podvig:2005:IBD

[Pod05] Pavel Podvig. If It’s broke, don’t fix it. Bulletin of the Atomic Scientists, 61(4):21–22, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/4/21.

Podvig:2016:BLB

[Pod16] Pavel Podvig. Blurring the line between nuclear and nonnuclear weapons: Increasing the risk of accidental nuclear war? Bulletin of the Atomic Scientists, 72(3):145–149, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Podvig:2024:PNT

[Pod24] Pavel Podvig. Preserving the nuclear test ban after Russia revoked its CTBT ratification. Bulletin of the Atomic Scientists, 80(2):75–80, 2024. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2024.2314431.

Pauling:1949:MSC

[POF49] Linus Pauling, Sir John Boyd Orr, and Enrico Fermi. More scientists comment on Soviet bomb. Bulletin of the Atomic Scientists, 5(11):322–324, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pogson:1954:BBL

[Pog54] Robert Pogson. Books: The Limits of the Earth, by Fairfield Osborn. Bulletin of the Atomic Scientists, 10(4):140, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1946:FFS

[Pol46] Michael Polanyi. The foundations of freedom in science. Bulletin of the Atomic Scientists, 2(11–12):6–7, December 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1949:OSP

[Pol49] Michael Polanyi. Ought science to be planned? two opposing views: The case for individualism. Bulletin of the Atomic Scientists, 5(1):19–20, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber49].

Polanyi:1950:FS

[Pol50a] Michael Polanyi. Freedom in science. Bulletin of the Atomic Scientists, 6(7):195–198, 224, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1950:SCF

[Pol50b] Michael Polanyi. Scientific convictions and the free society. Bulletin of the Atomic Scientists, 6(2):38–42, February 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1952:SBE

[Pol52] Michael Polanyi. Some British experiences. Bulletin of the Atomic Scientists, 8(7):223–228, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Polanyi:1953:PP

[Pol53] Michael Polanyi. Protests and problems. Bulletin of the Atomic Scientists, 9(9):322, 340, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1956:MM

[Pol56] Michael Polanyi. The magic of Marxism. Bulletin of the Atomic Scientists, 12(6):211–214, 232, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Fre57, Pol57a].

Polanyi:1957:LEM

[Pol57a] Michael Polanyi. Letter to the Editor: The magic of Marxism. Bulletin of the Atomic Scientists, 13(1):40, January 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pol56, Fre57].

Polanyi:1957:PVS

[Pol57b] Michael Polanyi. Part VIII: Science and philosophy: Passion and controversy in science. Bulletin of the Atomic Scientists, 13(4): 114–119, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1961:APS

[Pol61] John C. Polanyi. Armaments policies for the sixties. Bulletin of the Atomic Scientists, 17(10):403–406, 432–435, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1962:LEP

[Pol62a] John C. Polanyi. Letter to the Editor: Polanyi replies. Bulletin of the Atomic Scientists, 18(4):39–40, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kel62].

Polanyi:1962:LES

[Pol62b] John C. Polanyi. Letter to the Editor: Sealed monitors. Bulletin of the Atomic Scientists, 18(10):31, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polk:1966:SAI

[Pol66] William R. Polk. The scholar and the administrator in international affairs. Bulletin of the Atomic Scientists, 22(3):2–8, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polk:1967:MEA

[Pol67] William R. Polk. The Middle East: Analyzing social change. Bulletin of the Atomic Scientists, 23(1):12–19, January 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polach:1968:BBE

[Pol68] Jaroslav G. Polach. Books: Nuclear Energy in Poland During 1955–1960 and 1961–1963, edited by Josef Hurvic. Bulletin of the Atomic Scientists, 24(4):30–31, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polach:1969:NPE

[Pol69] Jaroslav G. Polach. Nuclear power in Europe at the crossroads. Bulletin of the Atomic Scientists, 25(8):15–18, 20, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polmar:1973:CNG

[Pol73] Norman Polmar. Communications: The numbers game. Bulletin of the Atomic Scientists, 29(5):3–4, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [EB73a, EB73b].

Polanyi:1976:NST

[Pol76] J. C. Polanyi. A nations’s science and technology. Bulletin of the Atomic Scientists, 32(3):8–12, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1979:CE

[Pol79a] John Polanyi. Cyrus Eaton 1883–1979. Bulletin of the Atomic Scientists, 35(8):7, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pollack:1979:LCM

[Pol79b] Jonathan D. Pollack. The logic of Chinese military strategy. Bulletin of the Atomic Scientists, 35(1):22–33, January 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1980:DNW

[Pol80] John C. Polanyi. The dangers of nuclear war. Bulletin of the Atomic Scientists, 36(1):6–10, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polachek:1988:LFE

[Pol88] Harry Polachek. Letter: Fatal errors. Bulletin of the Atomic Scientists, 44(10):48, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polanyi:1991:NWO

[Pol91a] John C. Polanyi. New world order: Collective will or law of the jungle. Bulletin of the Atomic Scientists, 47(5):23–24, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polner:1991:BRT

[Pol91b] Murray Polner. Book review: Not in their name: Uncommon Martyrs: THe Plowshares Movement and the Catholic Left, by Fred A. Wilcox. Bulletin of the Atomic Scientists, 47(7):42–43, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Poliski:1993:BRE

[Pol93] Iris Poliski. Book review: Exposing Americans: American Ground Zero, by Carole Gallagher. Bulletin of the Atomic Scientists, 49(5):50–51, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Polmar:2001:UI

[Pol01] Norman Polmar. Underwater intrigue. Bulletin of the Atomic Scientists, 57(5):66–68, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/5/66.

Pollack:2009:ESD

[Pol09a] Joshua Pollack. Emerging strategic dilemmas in U.S.–Chinese relations. Bulletin of the Atomic Scientists, 65(4):53–63, July 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/4/53.

Pollack:2009:ECP

[Pol09b] Joshua Pollack. Evaluating conventional prompt global strike. Bulletin of the Atomic Scientists, 65(1):13–20, January/February 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/65/1/13.

Pomerance:1961:POD

[Pom61] Josephin W. Pomerance. Public opinion and disarmament. Bulletin of the Atomic Scientists, 17(4):149–152, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pomerance:1965:MUC

[Pom65] Josephine W. Pomerance. Meeting U.N. crises: Let us seek basic solutions. Bulletin of the Atomic Scientists, 21(5):31–32, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pomerance:1966:BBC

[Pom66] Josephine W. Pomerance. Books: Crisis in the Congo, by Ernest Lefever. Bulletin of the Atomic Scientists, 22(1):41, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pomerance:1967:BBP

[Pom67] Josephine W. Pomerance. Books: International Peace Observation, by David W. Wainhouse and Associates. Bulletin of the Atomic Scientists, 23(1):29–30, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pomerance:1977:ATB

[Pom77] Jo Pomerance. The anti-test-ban coalition. Bulletin of the Atomic Scientists, 33(1):51–54, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pomerance:1979:CTB

[Pom79] Jo Pomerance. The comprehensive test ban at last? Bulletin of the Atomic Scientists, 35(7):9–10, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ponomarev:1960:PPC

[Pon60] Boris N. Ponomarev. [Principle of peaceful coexistence]. Bulletin of the Atomic Scientists, 16(7):288, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pontecorvo:1965:LES

[Pon65] Giulio Pontecorvo. Letter to the Editor: From the sea. Bulletin of the Atomic Scientists, 21(5):33–34, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pontecorvo:1967:BBC

[Pon67a] Giulio Pontecorvo. Books: The Common Wealth in Ocean Fisheries, by Francis T. Christy, Jr., and Anthony Scott. Bulletin of the Atomic Scientists, 23(6):55–56, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pontecorvo:1967:LS

[Pon67b] Giulio Pontecorvo. The law of the sea. Bulletin of the Atomic Scientists, 23(4):46–48, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PAS:1982:DPN

[Pon82] Pontifical Academy of Sciences. Declaration on prevention of nuclear war. Bulletin of the Atomic Scientists, 38(10):4–5, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Presented to His Holiness the Pope on 24 September 1982.

Poneman:2017:CAN

[Pon17] Daniel B. Poneman. The case for American nuclear leadership. Bulletin of the Atomic Scientists, 73(1):44–47, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2016.1264211.

PPXII:1952:SCC

[Pop52] Pope Pius XII. Science and the Catholic Church: Two documents. I: “Humani Generis”: Papal Encyclical, August 12, 1950. II: Theology and Modern Science: Pope Pius XII on the Harmony of the Work of God. Bulletin of the Atomic Scientists, 8 (5):142–146, 165, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PJXXIII:1964:PT

[Pop64] Pope John XXIII. Pacem in terris, 1963. Bulletin of the Atomic Scientists, 20(1):2, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PJPII:1981:PHS

[Pop81] Pope John Paul II. Perspective from Hiroshima: Science and conscience. Bulletin of the Atomic Scientists, 37(4):7–8, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Popova:1993:PRN

[Pop93] Lydia Popova. Perspective: Russia’s nuclear elite on rampage. Bulletin of the Atomic Scientists, 49(3):14–15, 47, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bro93].

Popescu:2019:ETA

[Pop19] Saskia Popescu. The existential threat of antimicrobial resistance. Bulletin of the Atomic Scientists, 75(6):286–289, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1680053.

Porter:1951:EPP

[Por51] Paul R. Porter. ECA: Plans and prospects. Bulletin of the Atomic Scientists, 7(7–8):220–222, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Porter:1971:CCB

[Por71] William W. Porter II. Communications: Competition between ecosystems. Bulletin of the Atomic Scientists, 27(3):3, 37, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Porter:1973:CEA

[Por73] William W. Porter II. Communications: Energy in America. Bulletin of the Atomic Scientists, 29(2):3–4, February 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Porter:1974:LEO

[Por74] William W. Porter II. Letter: The EPA and oil. Bulletin of the Atomic Scientists, 30(10):3, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tra73].

Porro:1982:PWB

[Por82] Jeffrey D. Porro. The policy war: Brodie vs. Kahn. Bulletin of the Atomic Scientists, 38(6):16–19, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Porras:2019:ASW

[Por19] Daniel Porras. Anti-satellite warfare and the case for an alternative draft treaty for space security. Bulletin of the Atomic Scientists, 75(4):142–147, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1628470.

Post:1971:FPU

[Pos71] Richard F. Post. Fusion power: The uncertain certainty. Bulletin of the Atomic Scientists, 27(8):42–48, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Postel:1979:BRB

[Pos79] Theodore Postel. Book reviews: The Fight Over Nuclear Energy, by Fred H. Schmidt and David Bodansky; Energy or Extinction, by Fred Hoyle. Bulletin of the Atomic Scientists, 35(1):44–45, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Postol:1985:SCN

[Pos85] Theodore A. Postol. Strategic confusion — with or without nuclear winter. Bulletin of the Atomic Scientists, 41(2):14–17, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Posen:1987:ARS

[Pos87] Barry R. Posen. Arms race at sea: U.S. maritime strategy: a dangerous game. Bulletin of the Atomic Scientists, 43(7):24–28, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Postol:2000:NMD

[Pos00] Theodore A. Postol. National missile defense: The target is Russia. Bulletin of the Atomic Scientists, 56(2):30–35, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/2/30.

Posen:2021:NTD

[Pos21] Barry R. Posen. A new transatlantic division of labor could save billions every year! Bulletin of the Atomic Scientists, 77 (5):239–243, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1964223.

Post:2024:FPU

[Pos24] Richard F. Post. Fusion power: The uncertain certainty. Bulletin of the Atomic Scientists, 80(6):385–393, 2024. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2024.2419768.

Potter:1949:LED

[Pot49] Zenas L. Potter. Letter to the Editor: Danger of the atomic bomb. Bulletin of the Atomic Scientists, 5(4):127–128, April 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Potts:1953:NPAa

[Pot53a] Ramsay D. Potts, Jr. National policy and air defense. Bulletin of the Atomic Scientists, 9(7):253–255, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Potts:1953:NPAb

[Pot53b] Ramsay D. Potts, Jr. National policy and air defense. Bulletin of the Atomic Scientists, 9(7):253–255, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Potter:1970:PND

[Pot70] Neal Potter. Part 4: The next decades: The Antarctic: Any economic future? Bulletin of the Atomic Scientists, 26(10):94– 99, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Potter:1997:PUS

[Pot97] William C. Potter. Perspective: Unsafe at any size. Bulletin of the Atomic Scientists, 53(3):25–27, 61, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Poulose:1978:CIA

[Pou78] T. T. Poulose. Commentary: From India: Atomic colonialism. Bulletin of the Atomic Scientists, 34(8):58–60, October 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powers:1947:OSF

[Pow47] Philip N. Powers. The organization for science in the Federal Government: A report on the existing government agencies for coordinating research. Bulletin of the Atomic Scientists, 3(4–5): 122–123, 126, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1959:IES

[Pow59] C. F. Powell. International exchange of scientists and students. Bulletin of the Atomic Scientists, 15(6):234–235, June 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powers:1974:LOT

[Pow74] Thomas Powers. Letter: The other targets. Bulletin of the Atomic Scientists, 30(9):3, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor74].

Powell:1981:HCH

[Pow81] John W. Powell. A hidden chapter in history. Bulletin of the Atomic Scientists, 37(8):44–52, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1982:SUS

[Pow82] David E. Powell. The Soviet Union: Social trends and social problems. Bulletin of the Atomic Scientists, 38(9):19–25, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1983:NPJ

[Pow83] John W. Powell. Nuclear power in Japan. Bulletin of the Atomic Scientists, 39(5):33–39, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powe:1986:ELF

[Pow86] Scot Powe. Espionage, leaks, and the First Amendment. Bulletin of the Atomic Scientists, 42(6):8–10, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1988:BRB

[Pow88] John Powell. Book review: The Genius of China, by Robert Temple. Bulletin of the Atomic Scientists, 44(1):52–53, January/ February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1989:PWT

[Pow89a] John Powell. Perspective: What the Third World can handle. Bulletin of the Atomic Scientists, 45(4):8, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1989:BRB

[Pow89b] John W. Powell. Book review: Edgar Snow, by John M. Hamilton. Bulletin of the Atomic Scientists, 45(5):41–42, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See omission [The89e].

Powell:1990:BRT

[Pow90a] John W. Powell. Book review: Ties that bind Entangling Alliances: How the Third World Shapes Our Lives, by John Maxwell Hamilton. Bulletin of the Atomic Scientists, 46(8):39–40, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powell:1990:LGM

[Pow90b] John W. Powell. Letter: Give the military golden parachutes. Bulletin of the Atomic Scientists, 46(7):46–47, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano90t].

Powell:1990:LIB

[Pow90c] John W. Powell. Letter: Ideological baggage. Bulletin of the Atomic Scientists, 46(1):46–47, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SH89b].

Powell:1991:LLM

[Pow91a] Joel W. Powell. Letter: Lost missiles found. Bulletin of the Atomic Scientists, 47(5):46–47, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [FP90c].

Powell:1991:BRB

[Pow91b] John W. Powell. Book review: Building blocks: Technology in World Civilization, by Arnold Pacey. Bulletin of the Atomic Scientists, 47(5):39, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powers:1993:HWS

[Pow93] Thomas Powers. Heisenberg’s war: the secret history of the German bomb. Alfred A. Knopf, New York, NY, USA, 1993. ISBN 0-394-51411-4. xi + 607 pp. LCCN QC16.H35 P69 1993. US$27.50, CAN$34.50.

Powaski:1998:NJM

[Pow98] Ronald E. Powaski. NATO: Joining the march of folly. Bulletin of the Atomic Scientists, 54(1):18–22, January/February 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Powaski:2001:AO

[Pow01] Ronald E. Powaski. An army of its own. Bulletin of the Atomic Scientists, 57(3):25–27, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/3/25.

Powaski:2004:BNH

[Pow04] Ronald E. Powaski. Bush’s nuclear hypocrisy. Bulletin of the Atomic Scientists, 60(1):24–25, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/1/24.

Pendred:1963:LEP

[PP63] Keith J. Pendred and C. Northcote Parkinson. Letter to the Editor: Pendred’s Law. Bulletin of the Atomic Scientists, 19 (3):30, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Par62].

Perlo:1983:CCR

[PP83] Victor Perlo and Christopher Paine. Commentary: On ‘Curtain rises on the European nuclear theater’. Bulletin of the Atomic Scientists, 39(10):56, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pai83c].

Park:2022:MLI

[PPE22] Sulgiye Park, Allison Puccioni, and Rodney C. Ewing. Machine learning improves satellite imagery analysis of North Korean nuclear activity. Bulletin of the Atomic Scientists, 78(1):26–37, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.2014238.

Pauling:1962:LES

[PR62] Linus Pauling and Eugene Rabinowitch. Letter to the Editor: Scientists in politics. Bulletin of the Atomic Scientists, 18(10): 32, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Plotkin:1985:LDN

[PR85] Sheldon C. Plotkin and David Rose. Letter: Decommissioning nuclear power plants. Bulletin of the Atomic Scientists, 41(10):56– 57, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ros85].

Prins:2008:KP

[PR08] Gwyn Prins and Steve Rayner. The Kyoto protocol. Bulletin of the Atomic Scientists, 64(1):45–48, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/1/45.

Pillai:2014:BRP

[PR14] S. Rajendran Pillai and M. V. Ramana. Breeder reactors: a possible connection between metal corrosion and sodium leaks. Bulletin of the Atomic Scientists, 70(3):49–55, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1989:BBB

[Pra89a] John Prados. Bulletin: Booming business. Bulletin of the Atomic Scientists, 45(9):7, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1989:BCI

[Pra89b] John Prados. Bulletin: Collateral intelligence. Bulletin of the Atomic Scientists, 45(9):6–7, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1989:CNT

[Pra89c] John Prados. China’s “new thinking” on nuclear arms. Bulletin of the Atomic Scientists, 45(5):32–35, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1992:RCW

[Pra92] John Prados. Report: The Cold War: High-flying spies. Bulletin of the Atomic Scientists, 48(7):11–12, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1993:NML

[Pra93a] John Prados. The new man at Langley. Bulletin of the Atomic Scientists, 49(6):35, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1993:WC

[Pra93b] John Prados. Woolsey and the CIA. Bulletin of the Atomic Scientists, 49(6):33–38, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prasso:1995:CBB

[Pra95] Sheri Prasso. Cambodia: a $3 billion boondoggle. Bulletin of the Atomic Scientists, 51(2):36–40, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1996:NRH

[Pra96] John Prados. No reform here. Bulletin of the Atomic Scientists, 52(5):55–59, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sta96].

Prados:1997:YCI

[Pra97] John Prados. You call that intelligence? Bulletin of the Atomic Scientists, 53(2):20–21, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:1998:PCC

[Pra98] John Prados. Perspective: Come down from the clouds. Bulletin of the Atomic Scientists, 54(5):18–19, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prados:2002:UCI

[Pra02] John Prados. Understanding central intelligence. Bulletin of the Atomic Scientists, 58(2):64–65, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/64.

Prados:2003:NW

[Pra03a] John Prados. A necessary war? Bulletin of the Atomic Scientists, 59(3):26–33, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/3/26.

Prados:2003:SWS

[Pra03b] John Prados. “Slow-walked and stonewalled”. Bulletin of the Atomic Scientists, 59(2):28–37, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/2/28.

Prados:2003:WRR

[Pra03c] John Prados. Will the real revisionists please stand up? Bulletin of the Atomic Scientists, 59(5):??, September/October 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/5/21.

Prados:2004:BB

[Pra04a] John Prados. Blindsided or blind? Bulletin of the Atomic Scientists, 60(4):27–37, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/4/27.

Prados:2004:INE

[Pra04b] John Prados. Intelligence: No easy fix. Bulletin of the Atomic Scientists, 60(5):17–19, September/October 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/5/17.

Prados:2004:PD

[Pra04c] John Prados. The pros from Dover. Bulletin of the Atomic Scientists, 60(1):44–51, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/1/44.

Prados:2004:UIR

[Pra04d] John Prados. U.S. intelligence: Reality check. Bulletin of the Atomic Scientists, 60(3):15–17, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/15.

Prados:2005:BB

[Pra05a] John Prados. Blind in Baghdad. Bulletin of the Atomic Scientists, 61(1):18–20, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/1/18.

Prados:2005:GBU

[Pra05b] John Prados. Good, bad, and ugly. Bulletin of the Atomic Scientists, 61(4):67–69, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/4/67.

Prados:2005:HQC

[Pra05c] John Prados. How Qaddafi came clean. Bulletin of the Atomic Scientists, 61(6):26–33, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/6/26.

Prados:2005:OH

[Pra05d] John Prados. Occupational hazards. Bulletin of the Atomic Scientists, 61(3):19–21, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/3/19.

Prados:2005:PPP

[Pra05e] John Prados. Pentagon power play. Bulletin of the Atomic Scientists, 61(2):21–23, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/61/2/21.

Prados:2006:BT

[Pra06a] John Prados. Brain trust. Bulletin of the Atomic Scientists, 62 (3):53–54, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/3/53.

Prados:2006:GP

[Pra06b] John Prados. Growing pains. Bulletin of the Atomic Scientists, 62(5):20–22, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/20.

Prados:2006:LV

[Pra06c] John Prados. The Long view. Bulletin of the Atomic Scientists, 62(1):22–23, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/1/22.

Pratt:2014:RR

[Pra14] Gill A. Pratt. Robot to the rescue. Bulletin of the Atomic Scientists, 70(1):63–69, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/70/1/63.full.pdf+html.

Prasad:2022:AFB

[Pra22] Eswar Prasad. After the fall: Bitcoin’s true legacy may be blockchain technology. Bulletin of the Atomic Scientists, 78 (4):187–190, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2087371.

PSRB:1948:SPP

[Pre48] President’s Scientific Research Board. Science and public policy. Bulletin of the Atomic Scientists, 4(1):23–29, 31, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PSAC:1958:AIO

[Pre58] President’s Science Advisory Committee. American “Introduction to outer space”. Bulletin of the Atomic Scientists, 14(5): 186–189, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prentice:1960:SIE

[Pre60] Edward S. Prentice. Significant issues in economic aid to newly developing countries. Bulletin of the Atomic Scientists, 16(8): 326–328, 333, October 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Preston:1962:BBP

[Pre62] Richard S. Preston. Books: Project Plowshare: The Development of the Peaceful Uses of Nuclear Explosions, by Ralph Sanders; Foreword by Willard F. Libby. Bulletin of the Atomic Scientists, 18(10):25–26, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Preston:1963:NRT

[Pre63] Richard Preston. News & reports: Test ban: Optimism on Project Vela. Bulletin of the Atomic Scientists, 19(6):33–36, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prescott:1975:BPO

[Pre75] James W. Prescott. Body pleasure and origins of violence. Bulletin of the Atomic Scientists, 31(9):10–20, November 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Koh76, Shi76, Lef76, DeM76, Bar76a].

Prescott:1976:VPR

[Pre76] James W. Prescott. Violence, pleasure, and religion. Bulletin of the Atomic Scientists, 32(3):62, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar76a].

Preiss:1980:COP

[Pre80] Kenneth Preiss. Commentary: Oil power. Bulletin of the Atomic Scientists, 36(2):64, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wil80c].

Press:1986:BRB

[Pre86] William H. Press. Book review: Star Warriors, by William J. Broad. Bulletin of the Atomic Scientists, 42(6):60, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Preer:1992:RCW

[Pre92] Robert Preer. Report: Conversion: Waging war for the tourist dollar. Bulletin of the Atomic Scientists, 48(3):7–8, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Press:1996:BRD

[Pre96] Eyal Press. Book review: The dismal facts: Nuclear Wastelands: A Global Guide to Nuclear Weapons Production and Its Health and Environmental Effects, edited by Arjun Makhijani, Howard Hu, and Katherine Yih. Bulletin of the Atomic Scientists, 52(2): 58–59, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pretorius:2017:BBB

[Pre17] Joelien Pretorius. Ban the bomb by … banning the bomb? Bulletin of the Atomic Scientists, 73(3):201–203, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1315121.

Price:1947:DNS

[Pri47] Don K. Price. The deficiencies of the National Science Foundation Bill. Bulletin of the Atomic Scientists, 3(10):291–294, 310, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pringsheim:1952:SBE

[Pri52] E. A. Pringsheim. Some British experiences. Bulletin of the Atomic Scientists, 8(7):232, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Price:1953:PSC

[Pri53] M. Philips Price. The Parliamentary and Scientific Committee of Great Britain. Bulletin of the Atomic Scientists, 9(3):76–80, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dew53].

Priestley:1955:SNF

[Pri55] J. B. Priestley. Sir Nuclear Fission. Bulletin of the Atomic Scientists, 11(8):293–294, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Iri56].

Price:1956:AEC

[Pri56] Melvin Price. Atomic energy in Congress. Bulletin of the Atomic Scientists, 12(10):373–375, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1958:LD

[Pri58a] Charles C. Price. A look at disarmament. Bulletin of the Atomic Scientists, 14(6):229–231, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1958:LEA

[Pri58b] George W. Price. Letter to the Editor: After missiles and satellites, what? Bulletin of the Atomic Scientists, 14(4):161, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab57b, War59].

Price:1968:ESA

[Pri68] Don K. Price. Educating for the Scientific Age. Bulletin of the Atomic Scientists, 24(8):26–32, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1969:WSC

[Pri69] William J. Price. Washington scene: The case for agency research. Bulletin of the Atomic Scientists, 25(4):34–36, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1970:BBC

[Pri70] Elizabeth Price. Books: Behavior Control, by Perry London. Bulletin of the Atomic Scientists, 26(7):44, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1974:GEP

[Pri74] Charles C. Price. Guest editorial: Political action for a livable world. Bulletin of the Atomic Scientists, 30(4):43–45, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cut75] and response [Pri75a].

Price:1975:LR

[Pri75a] Charles C. Price. Letters: Response. Bulletin of the Atomic Scientists, 31(2):4, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pri74, Cut75].

Primack:1975:NRS

[Pri75b] Joel Primack. Nuclear-reactor safety — introduction to the issues. Bulletin of the Atomic Scientists, 31(7):15–19, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hal76b].

Primack:1975:SRS

[Pri75c] Joel Primack. [Sidebar:] reactor safety glossary. Bulletin of the Atomic Scientists, 31(7):17, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pritchard:1975:LC

[Pri75d] Philip Pritchard. Letters: A contribution. Bulletin of the Atomic Scientists, 31(1):2–3, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day74b, Day75m].

Price:1976:LUB

[Pri76a] Charles C. Price. Letter: U.S. has been stalling. Bulletin of the Atomic Scientists, 32(4):2, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Primack:1976:LCA

[Pri76b] Joel Primack. Letter: Cynical assumption. Bulletin of the Atomic Scientists, 32(2):2, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sal75b].

Prince:1977:LFS

[Pri77] David C. Prince. Letter: Full speed ahead. Bulletin of the Atomic Scientists, 33(5):6–7, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Prince:1980:CFU

[Pri80] David C. Prince. Commentary: First use: already on the books. Bulletin of the Atomic Scientists, 36(10):59, December 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1981:CW

[Pri81a] Charles C. Price. Commentary: Weaponeers. Bulletin of the Atomic Scientists, 37(8):60–61, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Primack:1981:HRS

[Pri81b] Joel Primack. Human rights in the Southern Cone. Bulletin of the Atomic Scientists, 37(2):24–29, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Price:1986:LSC

[Pri86a] Peggy Price. Letter: Super collider. Bulletin of the Atomic Scientists, 42(8):54, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cro86].

Priceman:1986:LEC

[Pri86b] Mark Priceman. Letter: Economics of the cold war. Bulletin of the Atomic Scientists, 42(10):51, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AG86].

Priceman:1989:LNW

[Pri89] Mark Priceman. Letter: No World War III to stop. Bulletin of the Atomic Scientists, 45(10):49–50, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Primack:2002:PPD

[Pri02] Joel Primack. Pelted by paint, downed by debris. Bulletin of the Atomic Scientists, 58(5):24–25, September/October 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/5/24.

Priebe:2003:ALO

[Pri03] Miranda Priebe. Airborne laser: Overweight and oh-so-late. Bulletin of the Atomic Scientists, 59(3):18–20, May/June 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/3/18.

Prieto:2007:SS

[Pri07] Daniel B. Prieto. Off-the-shelf security. Bulletin of the Atomic Scientists, 63(3):76, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/3/76.

Proxmire:1985:CMA

[Pro85] William Proxmire. Congress must act on proliferation. Bulletin of the Atomic Scientists, 41(3):32–34, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [LP85b].

Prudori:1985:LSS

[Pru85] Tony Prudori. Letter: Soviet scientists. Bulletin of the Atomic Scientists, 41(2):50–51, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei85c].

Pryce:1948:APW

[Pry48] M. H. L. Pryce. Atomic power: What are the prospects? Bulletin of the Atomic Scientists, 4(8):245–248, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Przeworski:1982:MPP

[Prz82] Adam Przeworski. Marshall Plan for Poland. Bulletin of the Atomic Scientists, 38(3):7–8, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [JP82].

Patterson:1972:CED

[PS72] Walter C. Patterson and T. W. Schultz. Communications: Ecosystem doom. Bulletin of the Atomic Scientists, 28(7):2, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch72c].

Phillips:2004:LGN

[PS04] Alan Phillips and Steven Starr. Let’s go no-LOW. Bulletin of the Atomic Scientists, 60(3):20–22, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/20.

Perry:1987:LTM

[PT87] Edward C. Perry and Roland M. Timerbaev. Letters: Timerbaev and McCloy–Zorin. Bulletin of the Atomic Scientists, 43(3):51– 52, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tim87].

Papazian:1970:CJ

[PTNR70] Haig P. Papazian, John Toigo, Lavina Bryant Noto, and George R. Rahnestock. Communications: On Jensenism. Bulletin of the Atomic Scientists, 26(7):42–43, September 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lew70f].

PICASC:1955:ECS

[Pub55] Public Information Committee, Atomic Scientists of Chicago. Edward Condon’s security clearance. Bulletin of the Atomic Scientists, 11(2):66, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PROMIT:1968:SS

[Pub68] Public Relations Office, MIT. MIT and the Selective Service. Bulletin of the Atomic Scientists, 24(3):35, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PCC:1971:SIS

[Pug71] Pugwash Continuing Committee. SALT and international security. Bulletin of the Atomic Scientists, 27(10):17–19, December 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PCC:1973:PRE

[Pug73] Pugwash Continuing Committee. Pugwash: Raison d’être. Bulletin of the Atomic Scientists, 29(2):22–29, February 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1976:DSD

[Pug76] Pugwash Council. Disarmament, security and development: Pugwash scientists assess the issues of peace and war. Bulletin of the Atomic Scientists, 32(10):28–33, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1977:PSC

[Pug77] Pugwash Council. Peace and security in a changing world. Bulletin of the Atomic Scientists, 33(10):33–39, December 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1978:CBS

[Pug78] Pugwash Council. Commentary: From Bulgaria: Special responsibilities of scientists. Bulletin of the Atomic Scientists, 34(9):56– 59, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1980:SRC

[Pug80] Pugwash Council. The second review conference. Bulletin of the Atomic Scientists, 36(4):40–41, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1981:CBC

[Pug81] Pugwash Council. Commentary from Banff, Canada: Pugwash 1981. Bulletin of the Atomic Scientists, 37(10):53–56, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1983:SPC

[Pug83] Pugwash Council. Statement from the Pugwash Council. Bulletin of the Atomic Scientists, 39(9):48–50, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1986:PCS

[Pug86] Pugwash Council. Pugwash on common security. Bulletin of the Atomic Scientists, 42(9):46–48, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

PC:1988:PDD

[Pug88] Pugwash Council. Pugwash Dagomys declaration: insuring the survival of civilization. Bulletin of the Atomic Scientists, 44(9): 11, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pugliese:2002:PBB

[Pug02] David Pugliese. Panama: Bombs on the beach. Bulletin of the Atomic Scientists, 58(4):55–60, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/4/55.

Pugliese:2005:MDN

[Pug05] David Pugliese. Missile defense: No thanks, neighbor. Bulletin of the Atomic Scientists, 61(4):14–16, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/4/14.

Pullinger:2004:BF

[Pul04] Stephen Pullinger. Blair’s folly. Bulletin of the Atomic Scientists, 60(2):44–53, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/2/44.

Putzell:1952:PI

[Put52] Edwin J. Putzell, Jr. The prospects for industry. Bulletin of the Atomic Scientists, 8(8):275–277, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Puzanov:1993:LFP

[Puz93a] Oleg Puzanov. Looking forward to the past. Bulletin of the Atomic Scientists, 49(8):47, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Puzanov:1993:QT

[Puz93b] Oleg Puzanov. Quiet tensions. Bulletin of the Atomic Scientists, 49(8):30–32, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Primack:1972:SPS

[PvH72] Joel Primack and Frank von Hippel. Scientists, politics and SST: a critical review. Bulletin of the Atomic Scientists, 28(4):24–30, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [vH72].

Primack:1974:NRS

[PvH74] Joel Primack and Frank von Hippel. Nuclear reactor safety: The origins and issues of a vital debate. Bulletin of the Atomic Scientists, 30(8):5–12, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [HvH75].

Poole:1976:FPP

[PW76a] Alan D. Poole and Robert H. Williams. Flower power: prospects for photosynthetic energy. Bulletin of the Atomic Scientists, 32 (5):48–58, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [PW76b] and comment [Whi76a].

Poole:1976:LFP

[PW76b] Alan D. Poole and Robert H. Williams. Letter: Flower power [correction]. Bulletin of the Atomic Scientists, 32(7):2, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [PW76a].

Panofsky:1996:LNQ

[PWS^{+}96] Wolfgang K. H. Panofsky, Kevin W. Wenzel, Kory W. Sylvester, Gary Cerefice, and Alex DeVolpi. Letters: No quick fix for plutonium threat. Bulletin of the Atomic Scientists, 52(1):3, 59– 61, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DeV95].

Pentony:1963:LES

[PWW63] DeVere Pentony, Urban Whitaker, and Marshall Windmiller. Letter to the Editor: Study continues. Bulletin of the Atomic Scientists, 19(1):33, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Plesch:1998:GOS

[PY98] Daniel Plesch and Stephen Young. Guest opinion: Senseless policy. Bulletin of the Atomic Scientists, 54(6):4, November/ December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Pyati:1974:LAP

[Pya74] S. Pyati. Letter: ‘atomic politicians’. Bulletin of the Atomic Scientists, 30(8):3, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom74].

Quaranta:1980:BRJ

[Qua80] A. Alberigi Quaranta. Book review: J. Williams, Editor — All in our Time. The reminiscences of Twelve Nuclear Pioneers, The Bulletin of Atomic Scientists. Il Nuovo Cimento A, 55(3): 374, February 1, 1980. CODEN NIFAAM. ISSN 0369-3546 (print), 1826-9869 (electronic). URL http://link.springer.

com/article/10.1007/BF02902469.

QJN:1952:US

[Que52] Queen Juliana of The Netherlands. The universality of the sciences. Bulletin of the Atomic Scientists, 8(6):172, 206, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Quester:1967:CNN

[Que67] George H. Quester. Comments: Is the Nuclear Nonproliferation Treaty enough? Bulletin of the Atomic Scientists, 23(9):35–37, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Quester:1969:INN

[Que69] George H. Quester. Israel and the nuclear non-proliferation treaty. Bulletin of the Atomic Scientists, 25(6):7–9, 44–45, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Quester:1970:ICB

[Que70a] George H. Quester. India contemplates the bomb. Bulletin of the Atomic Scientists, 26(1):13–16, 48, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Quester:1970:PPP

[Que70b] George H. Quester. Paris, Pretoria, Peking, … proliferation? Bulletin of the Atomic Scientists, 26(8):12–16, October 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Quester:1981:CEC

[Que81] George H. Quester. Commentary: The emperor’s cloths, the Kremlin’s armor. Bulletin of the Atomic Scientists, 37(6):7–8, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Quigg:1984:CTW

[Qui84] Catherine Thiel Quigg. Commentary: Tritium warning. Bulletin of the Atomic Scientists, 40(3):56–57, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [HQL84].

R:1983:CCN

[R.83] A. B. R. Commentary from Canada: No borders in a nuclear war. Bulletin of the Atomic Scientists, 39(4):40–41, April 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:DAM

[RA67] Eugene Rabinowitch and Ruth Adams, editors. Debate the antiballistic missile. Bulletin of the Atomic Scientists, Chicago, IL. USA, 1967. 172 pp. LCCN UG632 .R3. Production editor: Irene Macauley.

Rabinowitch:1946:ARC

[Rab46a] Eugene Rabinowitch. Acheson Report and Chicago Draft Convention. Bulletin of the Atomic Scientists, 1(12):12–14, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1946:BH

[Rab46b] Eugene Rabinowitch. Before Hiroshima. Bulletin of the Atomic Scientists, 1(10):1, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1946:EDS

[Rab46c] Eugene Rabinowitch. Editorial: Deadlock or slow progress? Bulletin of the Atomic Scientists, 2(7–8):1, 29, October 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1946:EFS

[Rab46d] Eugene Rabinowitch. Editorial: Freedom of scientific publication. Bulletin of the Atomic Scientists, 2(11–12):1, 32, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1946:EIC

[Rab46e] Eugene Rabinowitch. Editorial: International cooperation of scientists. Bulletin of the Atomic Scientists, 2(5–6):1, September 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1946:EWW

[Rab46f] Eugene Rabinowitch. Editorial: Which way for American science? Bulletin of the Atomic Scientists, 2(9–10):1, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1946:VIC

[Rab46g] Eugene Rabinowitch. A victory and an impending crisis. Bulletin of the Atomic Scientists, 2(3–4):1, 32, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:EAE

[Rab47a] Eugene Rabinowitch. Editorial: Atomic energy — credit and debit, 1946. Bulletin of the Atomic Scientists, 3(1):1, 32, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:EBA

[Rab47b] Eugene Rabinowitch. Editorial: Britain and the atomic bomb. Bulletin of the Atomic Scientists, 3(11):313–314, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:EC

[Rab47c] Eugene Rabinowitch. Editorial comment. Bulletin of the Atomic Scientists, 3(12):373, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:EIA

[Rab47d] Eugene Rabinowitch. Editorial: If the UN Atomic Energy Commission fails. Bulletin of the Atomic Scientists, 3(7):169–170, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ELC

[Rab47e] Eugene Rabinowitch. Editorial: Let’s have clear thinking. Bulletin of the Atomic Scientists, 3(6):137–138, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ESW

[Rab47f] Eugene Rabinowitch. Editorial: Scientists and world government. Bulletin of the Atomic Scientists, 3(12):345–346, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ESB

[Rab47g] Eugene Rabinowitch. Editorial: Survival before progress. Bulletin of the Atomic Scientists, 3(10):281–282, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:EAS

[Rab47h] Eugene Rabinowitch. Editorial: The atomic secrets. Bulletin of the Atomic Scientists, 3(2):33, 68, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:EDU

[Rab47i] Eugene Rabinowitch. Editorial: The decision to use the bomb. Bulletin of the Atomic Scientists, 3(2):33, 68, February 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ESA

[Rab47j] Eugene Rabinowitch. Editorial: The Soviet amendment and Mr. Gromyko’s speech. Bulletin of the Atomic Scientists, 3(3):69, 100, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ESP

[Rab47k] Eugene Rabinowitch. Editorial: The Soviet plan for atomic energy control. Bulletin of the Atomic Scientists, 3(8):201–202, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ETD

[Rab47l] Eugene Rabinowitch. Editorial: The Truman Doctrine and the bomb. Bulletin of the Atomic Scientists, 3(4–5):101, 136, April/ May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:ETY

[Rab47m] Eugene Rabinowitch. Editorial: Two years after Hiroshima. Bulletin of the Atomic Scientists, 3(9):233–234, September 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:AEN

[Rab47n] Eugene Rabinowitch. UN atomic energy news. Bulletin of the Atomic Scientists, 3(6):164–165, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1947:WM

[Rab47o] Eugene Rabinowitch. Working for a miracle. Bulletin of the Atomic Scientists, 3(12):350, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Comment on [Szi47d].

Rabi:1948:PCR

[Rab48a] I. I. Rabi. Publication is the chief responsibility. Bulletin of the Atomic Scientists, 4(3):73, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bri48a].

Rabinowitch:1948:ADR

[Rab48b] Eugene Rabinowitch. Are atoms Democratic or Republican? Bulletin of the Atomic Scientists, 4(6):161–162, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1948:BRB

[Rab48c] Eugene Rabinowitch. Book review: Scientist in Russia, by Eric Ashby. Bulletin of the Atomic Scientists, 4(2):61–62, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano48y] with author correction.

Rabinowitch:1948:CE

[Rab48d] Eugene Rabinowitch. Comment by the Editor. Bulletin of the Atomic Scientists, 4(4):104–105, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor48a, vL48c].

Rabinowitch:1948:EYC

[Rab48e] Eugene Rabinowitch. Editorial: A year of civilian control of atomic energy. Bulletin of the Atomic Scientists, 4(2):33, 64, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1948:EAE

[Rab48f] Eugene Rabinowitch. Editorial: Atomic energy and the General Assembly. Bulletin of the Atomic Scientists, 4(11):322–323, 334, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Zim48].

Rabinowitch:1948:EPA

[Rab48g] Eugene Rabinowitch. Editorial: Previews of Armageddon. Bulletin of the Atomic Scientists, 4(9):258–260, September 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1948:ESR

[Rab48h] Eugene Rabinowitch. Everyone shares the responsibility. Bulletin of the Atomic Scientists, 4(3):75, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bri48a].

Rabinowitch:1948:NW

[Rab48i] Eugene Rabinowitch. The narrow way out. Bulletin of the Atomic Scientists, 4(6):185–188, 190, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1949:AEC

[Rab49a] Eugene Rabinowitch. Atomic Energy Commission, Congress, and scientists. Bulletin of the Atomic Scientists, 5(12):325–326, 329, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1949:CAF

[Rab49b] Eugene Rabinowitch. The ‘cleansing’ of AEC fellowships. Bulletin of the Atomic Scientists, 5(6–7):161–162, June/July 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1949:EHI

[Rab49c] Eugene Rabinowitch. Editorial: How not to investigate the Atomic Energy Commission. Bulletin of the Atomic Scientists, 5 (8–9):207–208, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1949:FF

[Rab49d] Eugene Rabinowitch. Forewarned — but not forearmed. Bulletin of the Atomic Scientists, 5(10):273–275, 292, October 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1949:PGS

[Rab49e] Eugene Rabinowitch. The purge of genetics in the Soviet Union. Bulletin of the Atomic Scientists, 5(5):130, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1950:CDL

[Rab50a] Eugene Rabinowitch. Civil defense: The long-range view. Bulletin of the Atomic Scientists, 6(8–9):226–230, August/September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Rabinowitch:1950:EAW

[Rab50b] Eugene Rabinowitch. Editorial: Atomic weapons and the Korean War. Bulletin of the Atomic Scientists, 6(7):194, 217, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1950:EWS

[Rab50c] Eugene Rabinowitch. Editorial: It’s not what’s said, it’s who says it. Bulletin of the Atomic Scientists, 6(5):130–132, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1950:ETM

[Rab50d] Eugene Rabinowitch. Editorial: The three main points of Professor Joliot-Curie. Bulletin of the Atomic Scientists, 6(6):163–165, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1950:MMI

[Rab50e] Eugene Rabinowitch, editor. Minutes to midnight; the international control of atomic energy, volume 1 of Atomic science and education series. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1950. 128 pp. LCCN TK9145 .A843 no. 1.

Rabinowitch:1951:AST

[Rab51a] Eugene Rabinowitch. Atomic spy trials: Heretical afterthoughts. Bulletin of the Atomic Scientists, 7(5):139–142, 157, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:EFY

[Rab51b] Eugene Rabinowitch. Editorial: Five years after. Bulletin of the Atomic Scientists, 7(1):3–5, 12, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:EIM

[Rab51c] Eugene Rabinowitch. Editorial: International meetings. Bulletin of the Atomic Scientists, 7(10):318–319, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:ESW

[Rab51d] Eugene Rabinowitch. Editorial: Scientific womanpower. Bulletin of the Atomic Scientists, 7(2):34, 37, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:ESL

[Rab51e] Eugene Rabinowitch. Editorial: Scientists and loyalty. Bulletin of the Atomic Scientists, 7(12):354–355, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:ELS

[Rab51f] Eugene Rabinowitch. Editorial: The labors of Sisyphus. Bulletin of the Atomic Scientists, 7(10):290–291, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:EOR

[Rab51g] Eugene Rabinowitch. Editorial: The only real defense. Bulletin of the Atomic Scientists, 7(9):242–243, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:EWCa

[Rab51h] Eugene Rabinowitch. Editorial: We can’t play that game. Bulletin of the Atomic Scientists, 7(6):162–163, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:EWCb

[Rab51i] Eugene Rabinowitch. Editorial: Where credit is due. Bulletin of the Atomic Scientists, 7(12):355, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:EWW

[Rab51j] Eugene Rabinowitch. Editorial: World War III — or next week’s movie? Bulletin of the Atomic Scientists, 7(11):322, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:KMR

[Rab51k] Eugene Rabinowitch. Korea must be rebuilt — a task for the UN. Bulletin of the Atomic Scientists, 7(7–8):194–195, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:NN

[Rab51l] Eugene Rabinowitch. News and notes. Bulletin of the Atomic Scientists, 7(9):276–277, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1951:WPR

[Rab51m] Eugene Rabinowitch. World population and resources. Bulletin of the Atomic Scientists, 7(4):98, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:BAA

[Rab52a] Eugene Rabinowitch. A bill to amend the Atomic Energy Act of 1946, as amended. Bulletin of the Atomic Scientists, 8(1): 16, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:EHL

[Rab52b] Eugene Rabinowitch. Editorial: How to lose friends. Bulletin of the Atomic Scientists, 8(1):1–2, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bat52].

Rabinowitch:1952:ETY

[Rab52c] Eugene Rabinowitch. Editorial: Ten years after. Bulletin of the Atomic Scientists, 8(9):294–295, 314, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:HBS

[Rab52d] Eugene Rabinowitch. The “Hydrogen Bomb” story. Bulletin of the Atomic Scientists, 8(9):297–300, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:MSB

[Rab52e] Eugene Rabinowitch. The miracles of Soviet biology. Bulletin of the Atomic Scientists, 8(2):41, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:PSS

[Rab52f] Eugene Rabinowitch. “Planning” in Soviet science. Bulletin of the Atomic Scientists, 8(2):47, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:SBM

[Rab52g] Eugene Rabinowitch. Senator Brien McMahon [obituary]. Bulletin of the Atomic Scientists, 8(6):170, 201, August 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:UES

[Rab52h] Eugene Rabinowitch. USSR exchange of scientific information. Bulletin of the Atomic Scientists, 8(2):48–49, February 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:VSS

[Rab52i] Eugene Rabinowitch. V: Science and scientists in Russia. Bulletin of the Atomic Scientists, 8(3):74–78, March 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:BBN

[Rab53a] Eugene Rabinowitch. Books: The New Force, by Ralph Lapp. Bulletin of the Atomic Scientists, 9(8):316, October 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:BNY

[Rab53b] Eugene Rabinowitch. The Bulletin in the New Year. Bulletin of the Atomic Scientists, 9(1):3, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:EPC

[Rab53c] Eugene Rabinowitch. Editorial: Packing the courts of science? Bulletin of the Atomic Scientists, 9(4):98–99, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:ESI

[Rab53d] Eugene Rabinowitch. The exchange of scientific information with the Soviet Union. Bulletin of the Atomic Scientists, 9(1):13–15, 22, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:NW

[Rab53e] Eugene Rabinowitch. The narrowing way. Bulletin of the Atomic Scientists, 9(8):294–295, 298, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:RBO

[Rab53f] Eugene Rabinowitch. A review: Fortune’s own “Operation Candor”. Bulletin of the Atomic Scientists, 9(10):372–374, 382, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:SFD

[Rab53g] Eugene Rabinowitch. Science faces a double danger. Bulletin of the Atomic Scientists, 9(2):34–35, 42, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1953:SS

[Rab53h] Eugene Rabinowitch. Scientists as sleuths. Bulletin of the Atomic Scientists, 9(9):334–336, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EMM

[Rab54a] Eugene Rabinowitch. Editorial: Must millions march? Bulletin of the Atomic Scientists, 10(6):194–195, 238, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EPM

[Rab54b] Eugene Rabinowitch. Editorial: People must know. Bulletin of the Atomic Scientists, 10(10):370, 398, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EAC

[Rab54c] Eugene Rabinowitch. Editorial: The American Chemical Society and Madame Joliot-Curie. Bulletin of the Atomic Scientists, 10(3):66–67, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EHB

[Rab54d] Eugene Rabinowitch. Editorial: The hydrogen bomb and the great unsolved problems. Bulletin of the Atomic Scientists, 10 (5):146–147, 168, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EIP

[Rab54e] Eugene Rabinowitch. Editorial: The international pool. Bulletin of the Atomic Scientists, 10(4):114–115, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EWS

[Rab54f] Eugene Rabinowitch. Editorial: What is a security risk? Bulletin of the Atomic Scientists, 10(6):241, 256, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:EJE

[Rab54g] Eugene Rabinowitch. Europe: July 1954. I. Europe returns to normalcy. Bulletin of the Atomic Scientists, 10(8):319–322, 336, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:SP

[Rab54h] Eugene Rabinowitch. A speech for the President. Bulletin of the Atomic Scientists, 10(1):5–7, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1954:WSR

[Rab54i] Eugene Rabinowitch. What is a security risk? Bulletin of the Atomic Scientists, 10(7):241, 256, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1955:EEG

[Rab55a] Eugene Rabinowitch. Editorial: Eisenhower’s Geneva proposal. Bulletin of the Atomic Scientists, 11(7):234–235, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1955:EGG

[Rab55b] Eugene Rabinowitch. Editorial: Genetics in Geneva. Bulletin of the Atomic Scientists, 11(9):314–316, 343, November 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1955:LHB

[Rab55c] Eugene Rabinowitch. Living with H-bombs. Bulletin of the Atomic Scientists, 11(1):5–8, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:BTC

[Rab56a] Eugene Rabinowitch. The bomb test controversy. Bulletin of the Atomic Scientists, 12(9):322–323, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:EHB

[Rab56b] Eugene Rabinowitch. Editorial: H-bombs without fall-out. Bulletin of the Atomic Scientists, 12(7):224, 264, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:ELF

[Rab56c] Eugene Rabinowitch. Editorial: The lessons of a fateful month. Bulletin of the Atomic Scientists, 12(10):354–355, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:ECT

[Rab56d] Eugene Rabinowitch. Editor’s correction: [typos in editorials in the January and February 1956 issues of the Bulletin]. Bulletin of the Atomic Scientists, 12(3):92, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:ESI

[Rab56e] Eugene Rabinowitch. Europe — summer 1955: International cooperation of atomic scientists. Bulletin of the Atomic Scientists, 12(2):34–37, 61, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:LAE

[Rab56f] Eugene Rabinowitch. Letter to Art Editor. Bulletin of the Atomic Scientists, 12(2):57, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the introduction: “In my article Ten Years That Changed the World, I mentioned that in 1947, Dr. Teller ‘sketched the now familiar design of the Bulletin cover.’ ”.

Rabinowitch:1956:PIS

[Rab56g] Eugene Rabinowitch. Part III: Science and social responsibility: History’s challenge to scientists. Bulletin of the Atomic Scientists, 12(7):238–240, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:SFS

[Rab56h] Eugene Rabinowitch. Sir Francis Simon, July 2, 1893–October 31, 1956. Bulletin of the Atomic Scientists, 12(10):382, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:TYC

[Rab56i] Eugene Rabinowitch. Ten years that changed the world. Bulletin of the Atomic Scientists, 12(1):2–6, 32, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:TASc

[Rab56j] Eugene Rabinowitch. Tenth anniversary symposium: Science and the affairs of man: A last chance? Bulletin of the Atomic Scientists, 12(6):186–188, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:TASa

[Rab56k] Eugene Rabinowitch. Tenth anniversary symposium: Science and the affairs of man: Part I: Science and world politics. Bulletin of the Atomic Scientists, 12(4):98, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:TASb

[Rab56l] Eugene Rabinowitch. Tenth anniversary symposium: Science and the affairs of man: Part II: Science and military strategy. Bulletin of the Atomic Scientists, 12(5):138–139, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1956:WPN

[Rab56m] Eugene Rabinowitch. Walking the plank to nowhere. Bulletin of the Atomic Scientists, 12(8):282–283, 320, October 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1957:AD

[Rab57a] Eugene Rabinowitch. About disarmament. Bulletin of the Atomic Scientists, 13(8):277–282, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1957:AMS

[Rab57b] Eugene Rabinowitch. After missiles and satellites, what? Bulletin of the Atomic Scientists, 13(10):346–350, December 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [Pri58b, War59].

Rabinowitch:1957:APS

[Rab57c] Eugene Rabinowitch. After Pugwash: The Soviet reaction. Bulletin of the Atomic Scientists, 13(9):314–315, November 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kra57, Hoc57].

Rabinowitch:1957:EIN

[Rab57d] Eugene Rabinowitch. Editorial introduction: The nuclear weapon test ban. Bulletin of the Atomic Scientists, 13(6):201, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1957:FYD

[Rab57e] Eugene Rabinowitch. The first year of deterrence. Bulletin of the Atomic Scientists, 13(1):2–8, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1957:FM

[Rab57f] Eugene Rabinowitch. The frozen map. Bulletin of the Atomic Scientists, 13(6):208–211, 215, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1957:PCP

[Rab57g] Eugene Rabinowitch. The Pugwash Conference: Pugwash — history and outlook. Bulletin of the Atomic Scientists, 13(7): 243–248, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1957:RSS

[Rab57h] Eugene Rabinowitch. The role of science and scientists in public life. Bulletin of the Atomic Scientists, 13(3):80–81, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fin57c].

Rabinowitch:1958:AWM

[Rab58a] Eugene Rabinowitch. Accidental war, missiles, and world community. Bulletin of the Atomic Scientists, 14(6):202–204, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1958:FJC

[Rab58b] Eugene Rabinowitch. Frédéric Joliot Curie, 1900–1958. Bulletin of the Atomic Scientists, 14(9):399, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1958:NYT

[Rab58c] Eugene Rabinowitch. New Year’s thoughts. Bulletin of the Atomic Scientists, 14(1):2–6, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1958:NBT

[Rab58d] Eugene Rabinowitch. Nuclear bomb tests. Bulletin of the Atomic Scientists, 14(8):282–287, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1958:SBT

[Rab58e] Eugene Rabinowitch. Stop before turning. Bulletin of the Atomic Scientists, 14(7):250–255, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1958:TPC

[Rab58f] Eugene Rabinowitch. The third Pugwash conference. Bulletin of the Atomic Scientists, 14(9):338–340, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:BCWa

[Rab59a] Eugene Rabinowitch. The Berlin crisis: Will deterrence work? Bulletin of the Atomic Scientists, 15(4):151–154, April 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fle59] and reply [Rab59c].

Rabinowitch:1959:EN

[Rab59b] Eugene Rabinowitch. Editor’s note. Bulletin of the Atomic Scientists, 15(5):222, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Apology for the omission of the name of Academician G. Nadjakov of Bulgaria from the list of signers of the Vienna Declaration [Ano58l].

Rabinowitch:1959:ERB

[Rab59c] Eugene Rabinowitch. Editor’s reply: Berlin. Bulletin of the Atomic Scientists, 15(7):321, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab59a, Fle59].

Rabinowitch:1959:SQQ

[Rab59d] Eugene Rabinowitch. Status Quo with a Quid Pro Quo. Bulletin of the Atomic Scientists, 15(7):288–292, September 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pac59].

Rabinowitch:1959:FTF

[Rab59e] Eugene Rabinowitch. First things first. Bulletin of the Atomic Scientists, 15(9):361–365, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:IST

[Rab59f] Eugene Rabinowitch. The impact of science and technology. Bulletin of the Atomic Scientists, 15(10):410–411, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:ISA

[Rab59g] Eugene Rabinowitch. Integral science and atomized art. Bulletin of the Atomic Scientists, 15(2):64–67, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:LR

[Rab59h] Eugene Rabinowitch. Lysenko Redevivus. Bulletin of the Atomic Scientists, 15(6):271, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:BCWb

[Rab59i] Eugene Rabinowitch. On biological and chemical warfare. Bulletin of the Atomic Scientists, 15(8):337–339, October 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:RSA

[Rab59j] Eugene Rabinowitch. Responsibilities of scientists in the Atomic Age. Bulletin of the Atomic Scientists, 15(1):2–7, January 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:SPP

[Rab59k] Eugene Rabinowitch. Science and party platforms. Bulletin of the Atomic Scientists, 15(10):405–409, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:TPG

[Rab59l] Eugene Rabinowitch. A tax for peace by a group of friends. Bulletin of the Atomic Scientists, 15(10):412, December 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:TBT

[Rab59m] Eugene Rabinowitch. Test ban: Two views. Bulletin of the Atomic Scientists, 15(3):98–99, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:BDN

[Rab59n] Eugene Rabinowitch. [The biological dangers of nuclear weapons tests]. Bulletin of the Atomic Scientists, 15(6):259, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1959:WF

[Rab59o] Eugene Rabinowitch. We found out. Bulletin of the Atomic Scientists, 15(10):430, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Supplies missing author, Marya Mannes, and source, The Reporter October 1959, of poem On Limited Warfare on page 343 of October 1959 issue of The Bulletin.

Rabinowitch:1959:WS

[Rab59p] Eugene Rabinowitch. We’re sorry. Bulletin of the Atomic Scientists, 15(2):93, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Correction noting that [Bri58] is a condensation of a paper given at the Amherst Humanities Center for Liberal Education.

Rabinowitch:1960:AAD

[Rab60a] Eugene Rabinowitch. The Atomic Age doctrine: Common efforts for common aims. Bulletin of the Atomic Scientists, 16(7):258– 263, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Sta60a].

Rabinowitch:1960:BBF

[Rab60b] Eugene Rabinowitch. Books: The Fearful Choice: A Debate on Nuclear Policy, by Philip Toynbee. Bulletin of the Atomic Scientists, 16(5):188–189, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1960:DND

[Rab60c] Eugene Rabinowitch. The dawn of a new decade. Bulletin of the Atomic Scientists, 16(1):2–6, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The author notes that the hands on the ‘clock of doom’ on the cover of the Bulletin have been moved back a bit.

Rabinowitch:1960:DA

[Rab60d] Eugene Rabinowitch. Defenders or avengers? Bulletin of the Atomic Scientists, 16(9):355–358, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1960:ENC

[Rab60e] Eugene Rabinowitch. Editor’s note: [Cyrus Eaton’s role in funding of the Pugwash conferences]. Bulletin of the Atomic Scientists, 16(9):344, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1960:FG

[Rab60f] Eugene Rabinowitch. The failure at Geneva. Bulletin of the Atomic Scientists, 16(2):34–37, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1960:IBL

[Rab60g] Eugene Rabinowitch. [Introduction to How to Live with the Bomb — and Survive]. Bulletin of the Atomic Scientists, 16(2):58, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Szi60a].

Rabinowitch:1960:MCW

[Rab60h] Eugene Rabinowitch. More on chemical warfare: Summary of a paper by W. H. Summerson. Bulletin of the Atomic Scientists, 16(6):252, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:ACD

[Rab61a] Eugene Rabinowitch. Arms control and disarmament. Bulletin of the Atomic Scientists, 17(4):122, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:BBa

[Rab61b] Eugene Rabinowitch. Berlin and beyond. Bulletin of the Atomic Scientists, 17(7):258–262, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Gra61].

Rabinowitch:1961:DH

[Rab61c] Eugene Rabinowitch. Dag Hammarskjold. Bulletin of the Atomic Scientists, 17(8):315, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Short obituary after the plane crash in the Congo.

Rabinowitch:1961:ESM

[Rab61d] Eugene Rabinowitch. “Eppur si muove”. Bulletin of the Atomic Scientists, 17(1):2–7, 35, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:HF

[Rab61e] Eugene Rabinowitch. Hail and farewell. Bulletin of the Atomic Scientists, 17(3):86–90, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:I

[Rab61f] Eugene Rabinowitch. Introduction. Bulletin of the Atomic Scientists, 17(4):123–125, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:MHH

[Rab61g] Eugene Rabinowitch. Man and his habitat: Part I. Bulletin of the Atomic Scientists, 17(8):322, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:SAS

[Rab61h] Eugene Rabinowitch. Soviet Academy of Sciences. Bulletin of the Atomic Scientists, 17(10):441–442, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:SES

[Rab61i] Eugene Rabinowitch. Space exploration in the service of science. Bulletin of the Atomic Scientists, 17(5–6):170–171, May/ June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:SE

[Rab61j] Eugene Rabinowitch. Statement by the Editor. Bulletin of the Atomic Scientists, 17(2):73, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to inaccuracies in the article [All60g].

Rabinowitch:1961:SC

[Rab61k] Eugene Rabinowitch. The Stowe Conferences. Bulletin of the Atomic Scientists, 17(9):382–384, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1961:BBb

[Rab61l] Eugene Rabinowitch. To build or not to build? Bulletin of the Atomic Scientists, 17(9):354–355, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab62i].

Rabinowitch:1961:WS

[Rab61m] Eugene Rabinowitch. We are sorry. Bulletin of the Atomic Scientists, 17(3):121, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Correction of misstatement on page 73 of the Bulletin: the journal is 15 years old, not 12.

Rabinowitch:1962:BCA

[Rab62a] Eugene Rabinowitch. B-70 in congress; atomic power; Soviet science. Bulletin of the Atomic Scientists, 18(5):46–47, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:CSP

[Rab62b] Eugene Rabinowitch. Conference on science and public affairs: Introduction. Bulletin of the Atomic Scientists, 18(1):15–16, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:EOD

[Rab62c] Eugene Rabinowitch. Editorial: An open door. Bulletin of the Atomic Scientists, 18(6):2–3, 24, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:ECE

[Rab62d] Eugene Rabinowitch. Events and comments: Escape into the past. Bulletin of the Atomic Scientists, 18(2):24, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:HOT

[Rab62e] Eugene Rabinowitch. Heros of our time. Bulletin of the Atomic Scientists, 18(5):25–26, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:MMP

[Rab62f] Eugene Rabinowitch. Man must prevail. Bulletin of the Atomic Scientists, 18(10):4–6, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:NYT

[Rab62g] Eugene Rabinowitch. New Year’s thoughts. Bulletin of the Atomic Scientists, 18(1):2–5, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:NIC

[Rab62h] Eugene Rabinowitch. Now it can be told. Bulletin of the Atomic Scientists, 18(8):25–28, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:RBB

[Rab62i] Eugene Rabinowitch. Response: To build or not to build. Bulletin of the Atomic Scientists, 18(3):39, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab61l].

Rabinowitch:1962:TOM

[Rab62j] Eugene Rabinowitch. A time for open minds. Bulletin of the Atomic Scientists, 18(7):2–3, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1962:TLU

[Rab62k] Eugene Rabinowitch. Together, let us explore the stars. Bulletin of the Atomic Scientists, 18(4):2, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:VT

[Rab63a] Eugene Rabinowitch. A. V. Topchiev: 1907–1962. Bulletin of the Atomic Scientists, 19(3):8–9, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:ACT

[Rab63b] Eugene Rabinowitch. After Cuba: Two lessons. Bulletin of the Atomic Scientists, 19(2):2–8, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [AG63].

Rabinowitch:1963:NYT

[Rab63c] Eugene Rabinowitch. New Year’s thoughts 1963. Bulletin of the Atomic Scientists, 19(1):2, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:PS

[Rab63d] Eugene Rabinowitch. Previews of space. Bulletin of the Atomic Scientists, 19(5):7, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:PCI

[Rab63e] Eugene Rabinowitch. Pugwash–Coswa: International conversations. Bulletin of the Atomic Scientists, 19(6):7–12, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:SRM

[Rab63f] Eugene Rabinowitch. Scientific revolution: Man’s new outlook. Bulletin of the Atomic Scientists, 19(7):15–18, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:SRB

[Rab63g] Eugene Rabinowitch. Scientific revolution: The beginning of world community. Bulletin of the Atomic Scientists, 19(10):14– 17, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:SRE

[Rab63h] Eugene Rabinowitch. Scientific revolution: The end of history. Bulletin of the Atomic Scientists, 19(9):9–12, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:SRN

[Rab63i] Eugene Rabinowitch. Scientific revolution: The new content of politics. Bulletin of the Atomic Scientists, 19(8):11–16, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1963:TBN

[Rab63j] Eugene Rabinowitch. Test ban now. Bulletin of the Atomic Scientists, 19(4):2, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1964:BBU

[Rab64a] Alex Rabinowitch. Books: The USSR and the Future, by Leonard Schapiro. Bulletin of the Atomic Scientists, 20(2):30, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1964:JFL

[Rab64b] Eugene Rabinowitch. James Franck 1882–1964, Leo Szilard 1898– 1964. Bulletin of the Atomic Scientists, 20(8):16–20, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1964:MG

[Rab64c] Eugene Rabinowitch. Morton Grodzins, 1917–1964. Bulletin of the Atomic Scientists, 20(5):15–16, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1964:NYT

[Rab64d] Eugene Rabinowitch. New Year’s thoughts 1964. Bulletin of the Atomic Scientists, 20(1):2, 18, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1964:SW

[Rab64e] Eugene Rabinowitch. Stop the world. Bulletin of the Atomic Scientists, 20(8):2–5, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1964:ICS

[Rab64f] Victor Rabinowitch. International cooperation in science: Unesco science and technology. Bulletin of the Atomic Scientists, 20 (9):31–34, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:AP

[Rab65a] Eugene Rabinowitch. About Pugwash. Bulletin of the Atomic Scientists, 21(4):9–15, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:AS

[Rab65b] Eugene Rabinowitch. Adlai E. Stevenson, 1900–1965. Bulletin of the Atomic Scientists, 21(7):2, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:BBS

[Rab65c] Eugene Rabinowitch. Books: Soviet Scientists in Red China, by Mikhail A. Klochko. Bulletin of the Atomic Scientists, 21(9): 29, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:ICSb

[Rab65d] Eugene Rabinowitch. International cooperation in science: Pugwash XIV. Bulletin of the Atomic Scientists, 21(7):41–45, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:NP

[Rab65e] Eugene Rabinowitch. The new perspectives. Bulletin of the Atomic Scientists, 21(9):2–3, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:NYT

[Rab65f] Eugene Rabinowitch. New Year’s thoughts 1965. Bulletin of the Atomic Scientists, 21(1):2–5, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1965:VFF

[Rab65g] Eugene Rabinowitch. Vietnam: Facts and fictions. Bulletin of the Atomic Scientists, 21(6):45–48, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GW65].

Rabinowitch:1965:ICSa

[Rab65h] Victor Rabinowitch. International cooperation in science: Biology in Euratom. Bulletin of the Atomic Scientists, 21(3):46–47, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1966:ECS

[Rab66a] Eugene Rabinowitch. The Editor comments: On space, basic research, and an International Science Foundation. Bulletin of the Atomic Scientists, 22(8):2–3, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1966:ECO

[Rab66b] Eugene Rabinowitch. The Editor comments: Open letter to Konrad Lorenz. Bulletin of the Atomic Scientists, 22(9):2–3, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Remarks on Lorentz’s 1966 book On Aggression. See comments [UM67].

Rabinowitch:1966:FOW

[Rab66c] Eugene Rabinowitch. Forum: Offense is the worst defense. Bulletin of the Atomic Scientists, 22(8):32–34, October 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1966:NYT

[Rab66d] Eugene Rabinowitch. New Year’s thoughts 1966. Bulletin of the Atomic Scientists, 22(1):3–7, 22, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1966:SC

[Rab66e] Eugene Rabinowitch. The second challenge. Bulletin of the Atomic Scientists, 22(5):25–27, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BS67].

Rabinowitch:1967:BBA

[Rab67a] Eugene Rabinowitch. Books: The Atom Bestrides the Nations, by V. S. Emelyanov. Bulletin of the Atomic Scientists, 23(1): 29, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:CSI

[Rab67b] Eugene Rabinowitch. Communications satellites: Introduction. Bulletin of the Atomic Scientists, 23(4):2–3, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:ECM

[Rab67c] Eugene Rabinowitch. The Editor comments: Missile gap and wheat gap. Bulletin of the Atomic Scientists, 23(2):2–3, 23, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Her67].

Rabinowitch:1967:ECF

[Rab67d] Eugene Rabinowitch. The Editor comments: On foreign aid. Bulletin of the Atomic Scientists, 23(6):2–4, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:ECP

[Rab67e] Eugene Rabinowitch. The Editor comments: On the pioneering generation. Bulletin of the Atomic Scientists, 23(9):2, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:IKI

[Rab67f] Eugene Rabinowitch. Igor Kurchatov, 1903–1960: An introduction. Bulletin of the Atomic Scientists, 23(10):8–10, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:NYT

[Rab67g] Eugene Rabinowitch. New Year’s thoughts 1967. Bulletin of the Atomic Scientists, 23(1):2–4, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:PXS

[Rab67h] Eugene Rabinowitch. Pugwash XVI [Sixteenth Pugwash Conference, Sopot, Poland, September 9–16, 1966]. Bulletin of the Atomic Scientists, 23(1):43, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1967:WSG

[Rab67i] Eugene Rabinowitch. What is sauce for the goose is sauce for the gander. Bulletin of the Atomic Scientists, 23(7):41–43, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1968:BBD

[Rab68a] Eugene Rabinowitch. Books: The Double Helix, by James Watson. Bulletin of the Atomic Scientists, 24(10):27–28, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1968:NYT

[Rab68b] Eugene Rabinowitch. New Year’s thoughts. Bulletin of the Atomic Scientists, 24(1):2–4, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1968:SMP

[Rab68c] Eugene Rabinowitch. The Sakharov Manifesto: Progress, peaceful coexistence, intellectual freedom. Bulletin of the Atomic Scientists, 24(9):2–7, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1968:SHC

[Rab68d] Eugene Rabinowitch. Science and the human condition. Bulletin of the Atomic Scientists, 24(8):23, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1968:SRA

[Rab68e] Eugene Rabinowitch. Student rebellion: The aimless revolution? Bulletin of the Atomic Scientists, 24(7):7–10, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mac68b, McG68].

Rabinowitch:1968:TCP

[Rab68f] Eugene Rabinowitch. The two challenges: The poor at home and abroad. Bulletin of the Atomic Scientists, 24(5):2–4, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1968:VHG

[Rab68g] Eugene Rabinowitch. The virus house: The German atomic bomb project: The German Atomic Bomb, by David Irving. Bulletin of the Atomic Scientists, 24(6):32–34, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabi:1969:ORS

[Rab69a] I. I. Rabi. Other reactions in the science community. Bulletin of the Atomic Scientists, 25(6):47, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:BBA

[Rab69b] Eugene Rabinowitch. Books: The American Challenge, by Jean-Jacques Servan-Schreiber, translated by Ronald Steel. Bulletin of the Atomic Scientists, 25(4):43–44, 47, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:CCS

[Rab69c] Eugene Rabinowitch. Conservation for conservation’s sake? Bulletin of the Atomic Scientists, 25(5):47–48, 56, May 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:AAW

[Rab69d] Eugene Rabinowitch. From Alamogordo to Apollo: Will man heed the lesson? Bulletin of the Atomic Scientists, 25(7):14– 16, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:NYT

[Rab69e] Eugene Rabinowitch. New Year’s thoughts, 1969: Prospects for progress and peace. Bulletin of the Atomic Scientists, 25(1):2–4, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:NMT

[Rab69f] Eugene Rabinowitch. NPT: Movement toward a viable world. Bulletin of the Atomic Scientists, 25(4):48, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:PCN

[Rab69g] Eugene Rabinowitch. Political criteria for non-political jobs? Bulletin of the Atomic Scientists, 25(6):2–3, 46, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:RA

[Rab69h] Eugene Rabinowitch. Reflections on Apollo 8. Bulletin of the Atomic Scientists, 25(3):2–3, 12, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:RSO

[Rab69i] Eugene Rabinowitch. Responsibility of scientists in our age. Bulletin of the Atomic Scientists, 25(9):2–3, November 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:SV

[Rab69j] Eugene Rabinowitch. Science and violence. Bulletin of the Atomic Scientists, 25(10):2–4, 44, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:SYR

[Rab69k] Eugene Rabinowitch. Scientists and youth in revolt. Bulletin of the Atomic Scientists, 25(4):2, 16–17, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SS69, Ame69].

Rabinowitch:1970:CFP

[Rab70a] Eugene Rabinowitch. Cecil Frank Powell. Bulletin of the Atomic Scientists, 26(5):48, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1970:JTT

[Rab70b] Eugene Rabinowitch. The Jensen thesis: Three comments: 3. Jensen vs Lewontin (a comment). Bulletin of the Atomic Scientists, 26(5):25–26, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1970:SCP

[Rab70c] Eugene Rabinowitch. On the Sochi Conference [19th Pugwash Conference on Science and World Affairs, 22–27 October 1969]. Bulletin of the Atomic Scientists, 26(4):18–20, April 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1970:TND

[Rab70d] Eugene Rabinowitch. On the threshold of a new decade: Thoughts for 1970: An editorial. Bulletin of the Atomic Scientists, 26(2):2–3, 46–59, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1970:TFY

[Rab70e] Eugene Rabinowitch. Twenty-five years later. Bulletin of the Atomic Scientists, 26(6):4–6, 34, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1971:BBC

[Rab71a] Eugene Rabinowitch. Books: The Coming War Between Russia and China, by Harrison Salisbury. Bulletin of the Atomic Scientists, 27(7):54–56, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1971:BBG

[Rab71b] Eugene Rabinowitch. Books: The Greening of America, by Charles A. Reich. Bulletin of the Atomic Scientists, 27(9):33– 37, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1971:MTU

[Rab71c] Eugene Rabinowitch. The mounting tide of unreason. Bulletin of the Atomic Scientists, 27(5):4–9, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [CSG71, Smi72b].

Rabinowitch:1971:RST

[Rab71d] Eugene Rabinowitch. The role of scientists: Thoughts for 1971. Bulletin of the Atomic Scientists, 27(1):2–4, January 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1972:BBR

[Rab72a] Eugene Rabinowitch. Books: The Rise and Fall of T. D. Lysenko, by Zhores A. Medvedev, translated by I. Michael Lerner; The Medvedev Papers — The Plight of Soviet Science Today, by Zhores A. Medvedev, translated by Vera Rich. Bulletin of the Atomic Scientists, 28(5):39–41, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1972:CMC

[Rab72b] Eugene Rabinowitch. Can man control his biological evolution? An introduction. Bulletin of the Atomic Scientists, 28(10):12, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1972:HTV

[Rab72c] Eugene Rabinowitch. Heretical thoughts. I. Vietnam: Politics vs. geography. Bulletin of the Atomic Scientists, 28(3):27–29, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Part II [Rab72d].

Rabinowitch:1972:HTI

[Rab72d] Eugene Rabinowitch. Heretical thoughts. II. Bengal: Balance of power revisited. Bulletin of the Atomic Scientists, 28(4):4–5, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Part I [Rab72c].

Rabinowitch:1972:MS

[Rab72e] Eugene Rabinowitch. The Moscow summit. Bulletin of the Atomic Scientists, 28(6):50, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1972:NPP

[Rab72f] Eugene Rabinowitch. Needed: A political program for the Technological Age. Bulletin of the Atomic Scientists, 28(7):11–15, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1972:TLD

[Rab72g] Eugene Rabinowitch. Thoughts for 1972 — living dangerously in the Age of Science. Bulletin of the Atomic Scientists, 28(1): 5–8, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Can72, Ger72].

Rabinowitch:1973:AVW

[Rab73a] Eugene Rabinowitch. After Vietnam, what? Bulletin of the Atomic Scientists, 29(6):13–15, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1973:BB

[Rab73b] Eugene Rabinowitch. Back into the bottle? Bulletin of the Atomic Scientists, 29(4):19–23, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1973:BBS

[Rab73c] Eugene Rabinowitch. Books: Primate Societies, by Hans Kummer. Bulletin of the Atomic Scientists, 29(3):56, March 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1973:BBM

[Rab73d] Eugene Rabinowitch. Books: Without Marx or Jesus: The New American Revolution Has Begun, by Jean-François Revel, translated by J. F. Bernard. Bulletin of the Atomic Scientists, 29(1): 46–48, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1973:BBB

[Rab73e] Eugene Rabinowitch. Books: World Without Borders, by Lester Brown. Bulletin of the Atomic Scientists, 29(6):46–48, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1973:CAA

[Rab73f] Eugene Rabinowitch. Challenges of the Atomic Age. Bulletin of the Atomic Scientists, 29(7):4–7, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Second of two articles that the author was working on at the time of his death on 15 May 1973.

Rabinowitch:1973:SYT

[Rab73g] Eugene Rabinowitch. Stopping, YES! but turning? — NO! Bulletin of the Atomic Scientists, 29(1):26–30, January 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabkin:1975:SJB

[Rab75] Yakov M. Rabkin. Social Jews: The bitter aftermath of emigration. Bulletin of the Atomic Scientists, 31(6):35–38, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rydberg:1980:CSN

[RAB^{+}80] Jan Rydberg, Dean Abrahamson, Wendy Barnaby, Thomas B. Johansson, and Peter Steen. Commentary: Sweden’s nuclear debate. Bulletin of the Atomic Scientists, 36(7):59–61, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AJSB79].

Rabinowitch:1985:ASC

[Rab85] Victor Rabinowitch. African science center thrives. Bulletin of the Atomic Scientists, 41(11):55–57, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1998:BRL

[Rab98] Marti Rabinowitch. Book review: Lethal legacy: Sowing the Dragon’s Teeth: Land Mines and the Global Legacy of War, by Philip C. Winslow. Bulletin of the Atomic Scientists, 54(1):65– 66, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:2005:FF

[Rab05a] Alexander Rabinowitch. Founder and father. Bulletin of the Atomic Scientists, 61(1):30–37, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/30.

Rabinowitch:2005:ACS

[Rab05b] Eugene Rabinowitch. Article commentary: a sustained reaction. Bulletin of the Atomic Scientists, 61(6):22–23, November/ December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/61/
6/22.

Rabinowitz:2016:NED

[Rab16] Or Rabinowitz. Nuclear energy and desalination in Israel. Bulletin of the Atomic Scientists, 72(1):32–38, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:2020:TYA

[Rab20] Eugene Rabinowitch. 1952: Ten years after. Bulletin of the Atomic Scientists, 76(6):353–355, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1847480.

Radin:1953:CPP

[Rad53] Alex Radin. The case for public power: Statement of American Public Power Association. Bulletin of the Atomic Scientists, 9 (4):134, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Radford:1983:BRB

[Rad83] Edward P. Radford. Book review: Radiation and Human Health, by John W. Gofman. Bulletin of the Atomic Scientists, 39(6):30– 31, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Radu:1992:CFS

[Rad92a] Michael Radu. Can Fujimori save Peru? Bulletin of the Atomic Scientists, 48(6):16–21, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sco92].

Radu:1992:PWD

[Rad92b] Michael Radu. Peru and the “War on Drugs”. Bulletin of the Atomic Scientists, 48(6):19, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Radkowsky:1997:LNF

[Rad97] Alvin Radkowsky. Letter: New fuel core, not a new reactor. Bulletin of the Atomic Scientists, 53(6):64–65, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fri97].

Rajaraman:2010:INA

[Raj10] Ramamurti Rajaraman. India’s nuclear arms control quandary. Bulletin of the Atomic Scientists, 66(2):27–36, March 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/2/27.

Rajaraman:2014:BWM

[Raj14] R. Rajaraman. Battlefield weapons and missile defense: Worrisome developments in nuclear South Asia. Bulletin of the Atomic Scientists, 70(2):68–74, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rajagopalan:2017:CPR

[Raj17] Rajesh Rajagopalan. China’s proper role in the global nuclear order. Bulletin of the Atomic Scientists, 73(2):133–134, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1288453.

Ramey:1964:RMG

[Ram64] James T. Ramey. The requirements merry-go-round: Must need precede development? Bulletin of the Atomic Scientists, 20 (9):12–15, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ramsey:1970:CCC

[Ram70] Norman F. Ramsey. The chemistry and cost of contamination: We need a pollution tax! Bulletin of the Atomic Scientists, 26 (4):3–5, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gol72b].

Ramsey:1972:CMC

[Ram72] Paul Ramsey. Can man control his biological evolution? — a symposium on genetic engineering: II. Genetic engineering. Bulletin of the Atomic Scientists, 28(10):14–17, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ramsey:1982:ABF

[Ram82] Norman F. Ramsey. August 1945: the B-29 flight logs. Bulletin of the Atomic Scientists, 38(10):33–35, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ramberg:1986:LIR

[Ram86a] Bennett Ramberg. Letter: Iraqi reactor. Bulletin of the Atomic Scientists, 42(7):65, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil86b].

Ramberg:1986:NPM

[Ram86b] Bennett Ramberg. Nuclear plants — military hostages? Bulletin of the Atomic Scientists, 42(3):17–21, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wil86b].

Ram:1988:MWE

[Ram88a] Bonnie J. Ram. Making warheads: Energy Department’s “good neighbor” policy. Bulletin of the Atomic Scientists, 44(1):35– 38, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ramberg:1988:BRB

[Ram88b] Bennett Ramberg. Book review: First Strike: The Exclusive Story of How Israel Foiled Iraq’s Attempt to Get the Bomb, by Shlomo Nakdimon. Bulletin of the Atomic Scientists, 44(5):48, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rambert:1997:LSR

[Ram97] Bennett Rambert. Letter: The semantics of reprocessing. Bulletin of the Atomic Scientists, 53(5):60, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hol97].

Ramberg:2004:DNM

[Ram04] Bennett Ramberg. Defusing the nuclear Middle East. Bulletin of the Atomic Scientists, 60(3):45–51, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/45.

Ramberg:2005:ANN

[Ram05] Bennett Ramberg. Avoiding the nuclear nightmare. Bulletin of the Atomic Scientists, 61(2):67–68, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/2/67.

Ramberg:2006:PP

[Ram06] Bennett Ramberg. Preemption paradox. Bulletin of the Atomic Scientists, 62(4):48–56, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/4/48.

Ramberg:2010:IQH

[Ram10] Bennett Ramberg. The Iranian quagmire: How to move forward: Position: The military option. Bulletin of the Atomic Scientists, 66(6):121–129, November/December 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/66/6/121.

Ramana:2011:NPP

[Ram11] M. V. Ramana. Nuclear power and the public. Bulletin of the Atomic Scientists, 67(4):43–51, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/67/4/43.

Ramana:2013:NPR

[Ram13a] M. V. Ramana. Nuclear policy responses to Fukushima: Exit, voice, and loyalty. Bulletin of the Atomic Scientists, 69(2): 66–76, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/2/66.

Ramana:2013:WIE

[Ram13b] M. V. Ramana. Why India’s electricity is likely to remain in short supply: The economics of nuclear power. Bulletin of the Atomic Scientists, 69(6):67–78, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/69/6/67.

Ramana:2016:COH

[Ram16] M. V. Ramana. The checkered operational history of hightemperature gas-cooled reactors. Bulletin of the Atomic Scientists, 72(3):171–179, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randers:1950:PAP

[Ran50] Gunnar Randers. Planning for atomic physics in Norway. Bulletin of the Atomic Scientists, 6(5):142, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randers:1953:DNA

[Ran53] Gunnar Randers. The Dutch–Norwegian atomic energy project. Bulletin of the Atomic Scientists, 9(10):369–371, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randers:1955:AED

[Ran55a] Gunnar Randers. [Atomic energy developments in the large underdeveloped areas of the Earth]. Bulletin of the Atomic Scientists, 11(6):220, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randers:1955:RG

[Ran55b] Gunnar Randers. The roads from Geneva. Bulletin of the Atomic Scientists, 11(10):358, 382, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randers:1959:IAE

[Ran59] Gunnar Randers. International Atomic Energy Agency. Bulletin of the Atomic Scientists, 15(4):163–167, April 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randers:1961:WAV

[Ran61] Gunnar Randers. What about Vienna? Bulletin of the Atomic Scientists, 17(1):20–23, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rankin:1982:CWO

[Ran82] William W. Rankin. Commentary: Where are our leaders? Bulletin of the Atomic Scientists, 38(6):70, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Randolph:1985:BRB

[Ran85] Robert H. Randolph. Book review: Contrasts: Soviet and American Thinkers Discuss the Future, by Wiktor Osiatyński, translated by Ewa Woydyło. Bulletin of the Atomic Scientists, 41 (10):50–51, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rao:1977:LTW

[Rao77] S. Ramachandra Rao. Letter: Third world fat cats. Bulletin of the Atomic Scientists, 33(3):4, 6, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sal76].

Rao:1987:IPR

[Rao87] R. V. R. Chandrasekhara Rao. India, Pakistan racing to be last. Bulletin of the Atomic Scientists, 43(9):32–34, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rapoport:1965:BBG

[Rap65a] Anatol Rapoport. Books: God and Golem, Inc.: A Comment on Certain Points where Cybernetics Impinges on Religion. Bulletin of the Atomic Scientists, 21(5):39–40, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rapoport:1965:SA

[Rap65b] Anatol Rapoport. The sources of anguish. Bulletin of the Atomic Scientists, 21(10):31–36, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bre65].

Rappert:2011:TMB

[Rap11] Brian Rappert. A teachable moment for biological weapons: The Seventh BWC Review Conference and the need for international cooperation in education. Bulletin of the Atomic Scientists, 67 (3):44–50, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/3/44.

Rashevsky:1955:BBE

[Ras55] N. Rashevsky. Books: Modern Experiments in Telepathy, by S. G. Soal and Frederick Bateman. Bulletin of the Atomic Scientists, 11(5):193–195, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raskin:1963:BBA

[Ras63] Marcus Raskin. Books: Deterrence, Arms Control, and Disarmament, by J. David Singer. Bulletin of the Atomic Scientists, 19(6):25, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rasmussen:1975:SSF

[Ras75] Norman C. Rasmussen. The safety study and its feedback. Bulletin of the Atomic Scientists, 31(7):25–28, September 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raskin:1980:CPD

[Ras80] Marcus G. Raskin. Commentary: the price of defense. Bulletin of the Atomic Scientists, 36(8):57–58, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raskin:1983:MZC

[Ras83] Marcus G. Raskin. The McCloy–Zorin correspondence. Bulletin of the Atomic Scientists, 39(2):34–36, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:2004:RD

[RAS04] Linda Rothstein, Catherine Auer, and Jonas Siegel. Rethinking doomsday. Bulletin of the Atomic Scientists, 60(6):36–73, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/6/36.

Rathjens:1958:DD

[Rat58] George W. Rathjens, Jr. Deterrence and defense. Bulletin of the Atomic Scientists, 14(6):225–228, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ratner:1962:LEW

[Rat62] Sidney Ratner. Letter to the Editor: Who is an expert? Bulletin of the Atomic Scientists, 18(4):40, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Fel62b].

Rathjens:1969:MAM

[Rat69] George W. Rathjens. Missiles and anti-missiles … six views: Is safeguard worth the risk? Bulletin of the Atomic Scientists, 25 (6):23–24, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rathjens:1971:AA

[Rat71a] George W. Rathjens. The ABCs of ABMs. Bulletin of the Atomic Scientists, 27(3):14–16, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rathjens:1971:EBA

[Rat71b] George W. Rathjens. An editorial: A breakthrough in arms control? Bulletin of the Atomic Scientists, 27(6):4–5, June 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rathjens:1972:SAA

[Rat72] George W. Rathjens. The SALT agreements: An appraisal. Bulletin of the Atomic Scientists, 28(9):8–10, November 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [You73].

Rathjens:1983:GBK

[Rat83] George W. Rathjens. George B. Kistiakowsky, 1900–1982. Bulletin of the Atomic Scientists, 39(4):2–3, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rathjens:1992:GIA

[Rat92] George Rathjens. The go-it-alone illusion. Bulletin of the Atomic Scientists, 48(4):32–33, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raushenbakh:1969:RAD

[Rau69] B. V. Raushenbakh. Reports: Automatic docking is Cosmos and its application to aviation safety and space stations. Bulletin of the Atomic Scientists, 25(1):40–42, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rauf:1991:BRB

[Rau91] Tariq Rauf. Book review: Baccanalia for the haves: Sleepwalking Through History: America in the Reagan Years, by Haynes Johnson. Bulletin of the Atomic Scientists, 47(7):45–46, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rauf:1992:RSU

[Rau92] Tariq Rauf. Report: Soviet Union: Cleaning up with a bang. Bulletin of the Atomic Scientists, 48(1):9, 47, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rauch:1996:LSF

[Rau96] Jerome S. Rauch. Letter: Stop fidgeting at the margin. Bulletin of the Atomic Scientists, 52(2):60–61, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo96c].

Rauf:1999:AC

[Rau99] Tariq Rauf. Accommodation, not confrontation. Bulletin of the Atomic Scientists, 55(1):14–16, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/14.

Ravetz:1965:FII

[Rav65] Jerome R. Ravetz. Forum: Irresponsibility and impossibility in politics — an intellectual’s defense. Bulletin of the Atomic Scientists, 21(3):31–33, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lev64, Lev65].

Ravenal:1983:NFU

[Rav83] Earl C. Ravenal. No first use: a view from the United States. Bulletin of the Atomic Scientists, 39(4):11–16, April 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SSR83].

Raven:1984:TWG

[Rav84] Peter H. Raven. Third World in the global future. Bulletin of the Atomic Scientists, 40(9):17–20, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ray:1980:CPH

[Ray80] Joel Ray. Citizens protest high power lines. Bulletin of the Atomic Scientists, 36(4):28–30, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SR81a].

Ray:1981:BRB

[Ray81] Joel Ray. Book review: The Geopolitics of Information: How Western Culture Dominates the World, by Anthony Smith. Bulletin of the Atomic Scientists, 37(8):40, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raymond:1989:LWH

[Ray89] Richard C. Raymond. Letter: Win hearts and minds. Bulletin of the Atomic Scientists, 45(4):52, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raymond:1990:LSS

[Ray90] Richard Raymond. Letter: 500 second-strike weapons are enough. Bulletin of the Atomic Scientists, 46(3):54–55, April 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH90a].

Russell:1965:LEF

[RB65] James W. Russell and Nathan M. Becker. Letter to the Editor: A flaw in pure science? Bulletin of the Atomic Scientists, 21 (7):28–29, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei65].

Rabinowitch:1967:JRO

[RB67] Eugene Rabinowitch and Hans Bethe. J. Robert Oppenheimer, 1904–1967. Bulletin of the Atomic Scientists, 23(8):2–6, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ring:1968:MS

[RC68] Gerald L. Ring and Eric W. Crawford. Missile submarines. Bulletin of the Atomic Scientists, 24(5):51, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ray:1970:TBC

[RC70] Edward Ray and Robert M. Cohen. “Tektite”: A blueprint for cooperative undersea scientific programs. Bulletin of the Atomic Scientists, 26(2):35–40, February 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rig70].

Roy-Chaudhury:1987:LAE

[RC87] Rahul Roy-Chaudhury. Letter: Aiming at empty silos. Bulletin of the Atomic Scientists, 43(6):60–61, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar87a].

Roy-Chaudhury:1988:LMD

[RC88] Rahul Roy-Chaudhury. Letter: Missiles don’t matter. Bulletin of the Atomic Scientists, 44(8):53–54, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kar88d].

Richards:1990:LAA

[RC90] Hugh T. Richards and John E. Carroll. Letter: Alzheimer’s, acid rain, and CFCs. Bulletin of the Atomic Scientists, 46(3):54, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Car89].

Rabkin:1975:LDV

[RD75] Yakov M. Rabkin and Samuel H. Day, Jr. Letters: Dramatization vs. responsibility. Bulletin of the Atomic Scientists, 31(7):3–4, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenbaum:1976:LSL

[RD76] David M. Rosenbaum and Samuel H. Day, Jr. Letter: Security with liberty. Bulletin of the Atomic Scientists, 32(7):6, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com76b].

Rowe:1977:LCM

[RD77] Lawrence A. Rowe and Sidney D. Drell. Letter: Counting missile shots. Bulletin of the Atomic Scientists, 33(8):2–3, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dre77].

Reppy:1980:BBT

[RD80] Judith Reppy and Harry Dean. Britain buys the Trident. Bulletin of the Atomic Scientists, 36(9):26–31, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Raffel:1991:STO

[RD91] James Raffel and Brian D’Agostino. START: Time for an old blueprint. Bulletin of the Atomic Scientists, 47(9):31–33, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Per92b].

Rabinowitch:1952:ISM

[RDS^{+}52] Eugene Rabinowitch, T. Harrison Davies, John A. Simpson, Jr., Michael Amrine, Peter Axel, and Mrs. Alan Simpson. An interview with Senator McMahon, Chairman, Joint Congressional Committee on Atomic Energy. Bulletin of the Atomic Scientists, 8(1):6–13, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rochlin:1985:LND

[RDS85] Gene I. Rochlin, Alex De Volpi, and Leonard S. Spector. Letter: Nonproliferation debate. Bulletin of the Atomic Scientists, 41(4): 61–64, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spe85a].

Russell:1963:AAW

[RE63] Bertrand Russell and Albert Einstein. An appeal for the abolition of war. In Grodzins and Rabinowitch [GR63], pages 539–541. LCCN D842 .B78. Reprint of [BEI^{+}55].

Rich:1968:PUS

[RE68] Alexander Rich and V. A. Engelhardt. A proposal from a U.S. and a Soviet scientist: Oceanic resources and developing nations. Bulletin of the Atomic Scientists, 24(2):2–3, February 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosner:2019:RAN

[RE19] Robert Rosner and Lynn Eden. Rebuilding an aging nuclear weapons complex: What should the United States do, and not do? An overview. Bulletin of the Atomic Scientists, 75(1):3– 8, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1555977.

Reagan:1968:CCB

[Rea68] Michael D. Reagan. Current comments: $17 billion in search of a policy. Bulletin of the Atomic Scientists, 24(4):33–36, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reader:1976:LPN

[Rea76] Mark Reader. Letter: The price of nuclear power. Bulletin of the Atomic Scientists, 32(4):3–4, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil75b].

Reader:1980:BRB

[Rea80] Mark Reader. Book reviews: Six Billion People: Demographic Dilemmas and World Politics, by George Tapinos and Phyllis T. Piotrow; World Futures: The Great Debate, edited by Christopher Freeman and Marie Jahoda. Bulletin of the Atomic Scientists, 36(2):48–49, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Record:1989:BRT

[Rec89] Jeffrey Record. Book review: Taking aim at the army: About Face, by Colonel David H. Hackworth. Bulletin of the Atomic Scientists, 45(8):37–38, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reding:1969:BBC

[Red69] Georges Reding. Books: The Cancer Ward, by Aleksandr I. Solzhenitsyn. Bulletin of the Atomic Scientists, 25(8):42, 44, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Redeker:1974:LO

[Red74] J. A. Redeker. Letter: An omission. Bulletin of the Atomic Scientists, 30(8):3, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano72h].

Reddy:1978:CEO

[Red78a] Amulya K. N. Reddy. Commentary: Energy options for the Third World. Bulletin of the Atomic Scientists, 34(10):54–55, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RL78b].

Reddy:1978:EOT

[Red78b] Amulya Kumar N. Reddy. Energy options for the Third World. Bulletin of the Atomic Scientists, 34(5):28–33, May 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reddy:1982:CRI

[Red82] Amulya Kumar N. Reddy. Commentary: Resolving India’s oil crisis. Bulletin of the Atomic Scientists, 38(4):47–49, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wil82b].

Rees:2007:GO

[Ree07] Martin Rees. Grounds for optimism. Bulletin of the Atomic Scientists, 63(1):32, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/32.

Rehman:2014:WAP

[Reh14] I. A. Rehman. A warning about Pakistan’s illusion of power. Bulletin of the Atomic Scientists, 70(6):73–78, November 2014. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reines:1950:PEU

[Rei50] Frederick Reines. Are the peaceful engineering users of atomic explosives? Bulletin of the Atomic Scientists, 6(6):171–172, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reines:1959:PNE

[Rei59] Frederick Reines. The peaceful nuclear explosion. Bulletin of the Atomic Scientists, 15(3):118–122, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reiner:1963:BBD

[Rei63] Thomas Reiner. Books: Unclassified Defense Contracts: Awards by County, State, and Metropolitan Areas of the United States Fiscal Year 1962, by Walter Isard and Gerald J. Karaska. Bulletin of the Atomic Scientists, 19(7):34–35, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reisman:1967:FAN

[Rei67] Davi Reisman. Forum: The answer is no. Bulletin of the Atomic Scientists, 23(6):64, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reinhold:1968:DS

[Rei68a] Robert Reinhold. A depression for science? Bulletin of the Atomic Scientists, 24(8):4–8, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reistrup:1968:BBH

[Rei68b] J. V. Reistrup. Books: Radiation Hazards in Uranium Mines. Bulletin of the Atomic Scientists, 24(4):32, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reissman:1972:BBL

[Rei72] Leonard Reissman. Books: The Lost Science of Man, by Ernest Becker. Bulletin of the Atomic Scientists, 28(1):46–47, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reissman:1974:BBI

[Rei74] Leonard Reissman. Books: The Intellectuals and the Powers, by Edward Shils. Bulletin of the Atomic Scientists, 30(1):46–47, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reiner:1978:BBD

[Rei78] Thomas A. Reiner. Books: International Disaster Relief: Toward a Responsive System, by Stephen Green; Reconstruction Following Disaster, edited by J. Eugene Haas, Robert W. Kates, and Martyn J. Bowden. Bulletin of the Atomic Scientists, 34(6):61– 62, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reiss:1983:G

[Rei83] Mitchell Reiss. GATT. Bulletin of the Atomic Scientists, 39(5): 53–56, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reilly:1991:LWN

[Rei91] Victor Reilly. Letter: Who needs tritium reactor? Bulletin of the Atomic Scientists, 47(10):44, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reilly:1993:LLT

[Rei93] Victor J. Reilly. Letter: Let tritium decay. Bulletin of the Atomic Scientists, 49(9):54–55, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reidy:2001:SC

[Rei01] Michael S. Reidy. Science in context. Bulletin of the Atomic Scientists, 57(2):75–76, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/75.

Reidy:2002:NEM

[Rei02] Michael S. Reidy. Never enough money. Bulletin of the Atomic Scientists, 58(4):65–66, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/65.

Remo:2015:DNE

[Rem15] John L. Remo. The dilemma of nuclear energy in space. Bulletin of the Atomic Scientists, 71(3):38–45, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Renner:1999:ACO

[Ren99] Michael Renner. Arms control orphans. Bulletin of the Atomic Scientists, 55(1):22–26, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/22.

Reppy:1983:BPR

[Rep83] Judith Reppy. On behalf of the public: the reports of Ruth Leger Sivard. Bulletin of the Atomic Scientists, 39(5):46–48, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reppy:1985:BRB

[Rep85a] Judith Reppy. Book review: The Arms Race: Social and Economic Consequences, by Hugh G. Mosley. Bulletin of the Atomic Scientists, 41(11):60–61, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reppy:1985:MRC

[Rep85b] Judith Reppy. Military R&D and the civilian economy. Bulletin of the Atomic Scientists, 41(9):10–14, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reppy:1986:BRB

[Rep86] Judith Reppy. Book reviews: A Reasonable Defense, by William W. Kaufman; Defense Choices: Greater Security with Fewer Dollars, by Jo Husbands. Bulletin of the Atomic Scientists, 42(7): 63, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reppy:1989:SPM

[Rep89] Judiciary Reppy. Science policy: More for the military. Bulletin of the Atomic Scientists, 45(1):46–48, January/February 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reparaz:2003:BMH

[Rep03] Adolfo Repáraz. Between MOX and a hard place. Bulletin of the Atomic Scientists, 59(5):46–51, September/October 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/5/46.

Reppy:2010:UNL

[Rep10] Judith Reppy. U.S. nuclear laboratories in a nuclear-zero world. Bulletin of the Atomic Scientists, 66(4):42–57, July/August 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/66/4/42.

Resnikoff:1976:LGE

[Res76] Marvin Resnikoff. Letter: The great equalizer. Bulletin of the Atomic Scientists, 32(7):6, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rethinaraj:1999:CS

[Ret99a] T. S. Gopi Rethinaraj. In the comfort of secrecy. Bulletin of the Atomic Scientists, 55(6):52–57, November/December 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/6/52.

Rethinaraj:1999:IF

[Ret99b] T. S. Gopi Rethinaraj. India: Frozen out. Bulletin of the Atomic Scientists, 55(3):15, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/3/15.

Revelle:1962:SRL

[Rev62] Roger Revelle. Some recent lessons. Bulletin of the Atomic Scientists, 18(1):20–24, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Revelle:1966:CMD

[Rev66] Roger Revelle. Can man domesticate himself? Bulletin of the Atomic Scientists, 22(2):2–7, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Revelle:1968:TAB

[Rev68] Roger Revelle. On technical assistance and bilateral aid. Bulletin of the Atomic Scientists, 24(3):17–19, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Revelle:1981:GSM

[Rev81] Roger Revelle. A good start, but miles to go. Bulletin of the Atomic Scientists, 37(9):27–32, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Revelle:1983:NLS

[Rev83] Roger Revelle. [Note on Law of the Sea]. Bulletin of the Atomic Scientists, 39(5):14–19, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Revelle:1992:RTW

[Rev92] Daniel Revelle. Report: Third World missiles: U.S. muscle misses mark. Bulletin of the Atomic Scientists, 48(9):10–11, 44, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Revkin:2021:BCP

[Rev21] Andrew Revkin. To build climate progress on time scales that matter, Biden should be Biden. Bulletin of the Atomic Scientists, 77(1):28–30, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1860327.

Rex:1971:GEN

[Rex71] Robert W. Rex. Geothermal energy — the neglected energy option. Bulletin of the Atomic Scientists, 27(8):52–56, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reynolds:2004:HT

[Rey04] Michael Reynolds. Homegrown terror. Bulletin of the Atomic Scientists, 60(6):48–57, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/6/48.

Rezek:2007:IBS

[Rez07a] John Rezek. Interview: Brice Smith. Bulletin of the Atomic Scientists, 63(6):22–27, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/6/22.

Rezek:2007:ISF

[Rez07b] John Rezek. Interview: Sergio Finardi. Bulletin of the Atomic Scientists, 63(5):20–25, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/5/20.

Rezek:2007:WLI

[Rez07c] John Rezek. Weapons lab: an interview with Matthew S. Meselson. Bulletin of the Atomic Scientists, 63(1):49–53, January/ February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/63/
1/49.

Rabinowitch:1955:AE

[RF55] Eugene Rabinowitch and James Franck. Albert Einstein. Bulletin of the Atomic Scientists, 11(6):202–203, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Romoser:1965:RSF

[RF65] George K. Romoser and Charles R. Foster. Reports: Safety first: The West German election. Bulletin of the Atomic Scientists, 21 (10):37–38, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1979:MAE

[RF79] Eugene Rabinowitch and James Franck. In memoriam [Albert Einstein]. Bulletin of the Atomic Scientists, 35(3):40, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Rosner:2021:NPS

[RF21] Robert Rosner and Sabrina Fields. Is nuclear power sustainable in a carbon-free world? The case of Sweden. Bulletin of the Atomic Scientists, 77(6):295–300, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1989196.

Rabinowitch:1949:ORE

[RG49] Eugene Rabinowitch and Hyman H. Goldsmith. To our readers: the Emergency Committee of Atomic Scientists and the Bulletin of the Atomic Scientists. Bulletin of the Atomic Scientists, 5(8–9): 261, August/September 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosner:2013:PRA

[RG13] Robert Rosner and Stephen M. Goldberg. A practical, regional approach to nuclear waste storage. Bulletin of the Atomic Scientists, 69(6):58–66, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/69/6/58.

Ray:1981:CMP

[RGA81] Roger Ray, Nathaniel A. Greenhouse, and Glenn H. Alcalay. Commentary: More on ‘Paradise Lost’. Bulletin of the Atomic Scientists, 37(8):62–63, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Alc81].

Rotblat:1995:LFP

[RGK95] Joseph Rotblat, Arttie Piar Gómez, and Milton Kirchman. Letters: False premises, wrong conclusions. Bulletin of the Atomic Scientists, 51(2):3, 69, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bai95].

Russell:1963:LEG

[RGR63] Bertrand Russell, Bentley Glass, and Eugene I. Rabinowitch. Letter to the Editor: Genetic effects. Bulletin of the Atomic Scientists, 19(6):31–32, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gla62c, Pau63, Gla63].

Rabinowitch:1965:RIB

[RH65] Victor Rabinowitch and Arthur D. Hasler. Reports: The International Biological Program. Bulletin of the Atomic Scientists, 21(8):32–34, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rossnagel:2012:LNS

[RH12] Alexander Rossnagel and Anja Hentschel. The legalities of a nuclear shutdown. Bulletin of the Atomic Scientists, 68(6):55–66, November/December 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/6/55.

Remo:2014:TSY

[RH14] John L. Remo and Hans J. Haubold. Threats from space: 20 years of progress. Bulletin of the Atomic Scientists, 70(4):85–93, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosner:2017:WRC

[RH17] Robert Rosner and Alex Hearn. What role could nuclear power play in limiting climate change? Bulletin of the Atomic Scientists, 73(1):2–6, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2016.1264203.

Rochlin:1978:WVR

[RHKK78] Gene I. Rochlin, Margery Held, Barbara G. Kaplan, and Lewis Kruger. West Valley: Remnant of the AEC. Bulletin of the Atomic Scientists, 34(1):17–23, 25–26, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rhoads:1946:MUA

[Rho46] C. P. Rhoads. The medical uses of atomic energy. Bulletin of the Atomic Scientists, 2(7–8):22–24, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rhodes:1990:PSS

[Rho90] Edward Rhodes. Perspective: Subversive submarines. Bulletin of the Atomic Scientists, 46(1):15, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rhodes:1995:DSM

[Rho95] Richard Rhodes. Dark Sun: The Making of the Hydrogen Bomb. Sloan technology series. Simon and Schuster, New York, NY, USA, 1995. ISBN 0-684-80400-X. 731 pp. LCCN UG1282.A8 R46 1995.

Rhodes:2007:WWS

[Rho07] Richard Rhodes. Why we should preserve the Manhattan Project. Bulletin of the Atomic Scientists, 63(3):34–43, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/63/3/34.

Rhodes:2009:RRB

[Rho09] Richard Rhodes. Richard Rhodes’ Reykjavik. Bulletin of the Atomic Scientists, 65(5):71–81, September 2009. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/65/5/71;   http://
bos.sagepub.com/content/65/5/71.abstract.

Rhodes:2013:GHI

[Rho13] Richard Rhodes. Guernica: Horror and inspiration. Bulletin of the Atomic Scientists, 69(6):19–25, November 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/6/19.

Rhodes:2015:WMP

[Rho15] Richard Rhodes. Why the Manhattan Project should be preserved. Bulletin of the Atomic Scientists, 71(6):4–10, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reford:1983:CCP

[RI83] Robert W. Reford and George Ignatieff. Commentary: On ‘The Canadian Prime Minister speaks’. Bulletin of the Atomic Scientists, 39(10):54–55, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tru83].

Rice:1952:RSS

[Ric52] Stuart A. Rice. Russian Science Symposium: VI: Statistics in the Soviet Union. Bulletin of the Atomic Scientists, 8(5):159–162, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richter:1953:FRV

[Ric53] Curt P. Richter. Free research versus design research. Bulletin of the Atomic Scientists, 9(9):323–324, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rich:1967:CED

[Ric67] Alexander Rich. Cooperative education in developing countries: Two programs. Bulletin of the Atomic Scientists, 23(9):43–45, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richard:1971:CGT

[Ric71] John Richard. Communications: Grand tour of the planets. Bulletin of the Atomic Scientists, 27(5):48, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richardson:1974:SAH

[Ric74] Robert A. Richardson. The selling of the atom: How Oak Ridge Associated Universities initiated a new kind of informational endeavor. Bulletin of the Atomic Scientists, 30(8):28–34, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rice:1975:PPU

[Ric75a] Stuart A. Rice. Playing politics with U.N. agencies. Bulletin of the Atomic Scientists, 31(7):7, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rice:1975:USD

[Ric75b] Stuart A. Rice. UNESCO’s self-destructive act. Bulletin of the Atomic Scientists, 31(2):6–7, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richter:1976:CSP

[Ric76] Maurice N. Richter, Jr. Chinese science policy: a comparative analysis. Bulletin of the Atomic Scientists, 32(3):13–16, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richmond:1977:LPH

[Ric77] Chester R. Richmond. Letter: Plutonium hazards. Bulletin of the Atomic Scientists, 33(4):6, 8, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eds76].

Richardson:1979:CNV

[Ric79] D’Arcy Richardson. Commentary: Newbomb vs. missliski: The great arm race. Bulletin of the Atomic Scientists, 35(3):81, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richter:1981:TFS

[Ric81] Roger Richter. Testimony from a former safeguards inspector. Bulletin of the Atomic Scientists, 37(8):29–31, October 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richter:1984:CRH

[Ric84] Paul H. Richter. Commentary: Reconsidering history. Bulletin of the Atomic Scientists, 40(10):32–33, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bal84a, Wei84].

Richelson:1986:OSN

[Ric86a] Jeffrey T. Richelson. Old surveillance, new interpretations. Bulletin of the Atomic Scientists, 42(2):18–23, February 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richelson:1986:UIS

[Ric86b] Jeffrey T. Richelson. U.S. intelligence and Soviet Star Wars. Bulletin of the Atomic Scientists, 42(5):12–14, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richelson:1989:CSM

[Ric89] Jeffrey Richelson. Commercial satellites: Military intelligence — SPOT is not enough. Bulletin of the Atomic Scientists, 45(7):26– 27, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richelson:1996:HFS

[Ric96] Jeffrey T. Richelson. High flyin’ spies. Bulletin of the Atomic Scientists, 52(5):48–54, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richelson:1998:DTP

[Ric98a] Jeffrey Richelson. Detection techniques: Past, present, and future. Bulletin of the Atomic Scientists, 54(6):54, November/ December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richelson:1998:VWM

[Ric98b] Jeffrey Richelson. Verification: The ways and means. Bulletin of the Atomic Scientists, 54(6):52–57, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Richelson:1999:UD

[Ric99] Jeffrey T. Richelson. Uncertain damage. Bulletin of the Atomic Scientists, 55(5):17–19, September/October 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/5/17.

Richelson:2000:DSS

[Ric00a] Jeffrey Richelson. Desperately seeking signals. Bulletin of the Atomic Scientists, 56(2):47–51, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/2/47.

Richelson:2000:SM

[Ric00b] Jeffrey T. Richelson. Shootin’ for the moon. Bulletin of the Atomic Scientists, 56(5):22–27, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://www.highbeam.com/doc/1G1-65230029.html.

Richelson:2001:LTS

[Ric01a] Jeffrey Richelson. Less than the sum of its parts. Bulletin of the Atomic Scientists, 57(1):70–71, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/70.

Richelson:2001:BB

[Ric01b] Jeffrey T. Richelson. Back to black. Bulletin of the Atomic Scientists, 57(3):22–24, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/3/22.

Richelson:2002:DNT

[Ric02] Jeffrey Richelson. Defusing nuclear terror. Bulletin of the Atomic Scientists, 58(2):38–43, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/2/38.

Richelson:2003:HB

[Ric03a] Jeffrey T. Richelson. Holding back. Bulletin of the Atomic Scientists, 59(6):26–32, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/6/26.

Richelson:2003:PD

[Ric03b] Jeffrey T. Richelson. Planning to deceive. Bulletin of the Atomic Scientists, 59(2):64–69, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/59/2/64.

Richelson:2003:SG

[Ric03c] Jeffrey T. Richelson. The satellite gap. Bulletin of the Atomic Scientists, 59(1):48–54, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/48.

Richelson:2005:SS

[Ric05] Jeffrey T. Richelson. Satellite in the shadows. Bulletin of the Atomic Scientists, 61(3):26–33, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/3/26.

Richelson:2006:WWW

[Ric06] Jeffrey T. Richelson. The whole world is watching. Bulletin of the Atomic Scientists, 62(1):26–35, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/26.

Richardson:2009:SNT

[Ric09] Jeff Richardson. Shifting from a nuclear triad to a nuclear dyad. Bulletin of the Atomic Scientists, 65(5):33–42, September 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/65/5/33.

Richardson:2012:LHN

[Ric12] David Richardson. Lessons from Hiroshima and Nagasaki: The most exposed and most vulnerable. Bulletin of the Atomic Scientists, 68(3):29–35, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/68/3/29.

Ridenour:1946:SS

[Rid46] Louis N. Ridenour. Secrecy in science. Bulletin of the Atomic Scientists of Chicago, 1(6):3, 8, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ridenour:1947:MSA

[Rid47] Louis N. Ridenour. Military support of American science, a danger? Bulletin of the Atomic Scientists, 3(8):221–223, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ridenour:1950:PHB

[Rid50a] Louis Ridenour. The physics of the H-bomb. Bulletin of the Atomic Scientists, 6(4):105, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprint of a section of the author’s article in Scientific American, March 1950.

Ridenour:1950:HER

[Rid50b] Louis N. Ridenour. How effective are radioactive poisons in warfare? Bulletin of the Atomic Scientists, 6(7):199–202, 224, July 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ridenour:1951:SFG

[Rid51] Louis N. Ridenour. Scientists and the Federal Government. Bulletin of the Atomic Scientists, 7(2):35–37, February 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riesman:1958:AW

[Rie58] David Riesman. Abundance for what? Bulletin of the Atomic Scientists, 14(4):135–139, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riesman:1959:PPP

[Rie59] David Riesman. Private people and public policy. Bulletin of the Atomic Scientists, 15(5):203–208, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riehl:1960:MCB

[Rie60] Herbert Riehl. Meteorological conditions and BW. Bulletin of the Atomic Scientists, 16(6):253, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riggs:1965:WIC

[Rig65] F. Behn Riggs. World information center. Bulletin of the Atomic Scientists, 21(1):34–35, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rigby:1970:CT

[Rig70] Richard N. Rigby, Jr. Communications: Tektite. Bulletin of the Atomic Scientists, 26(8):48, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RC70].

Riha:1965:PAR

[Rih65] Jeanne Riha. Peace in an Age of Revolution. Bulletin of the Atomic Scientists, 21(8):39–40, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ringuet:1952:FPU

[Rin52] M. Louis Leprince Ringuet. French physicists and U.S. visas. Bulletin of the Atomic Scientists, 8(7):247, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Ringo:1957:PDS

[Rin57] Roy Ringo. A practical defense system. Bulletin of the Atomic Scientists, 13(3):95–98, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ringer:1961:LEQ

[Rin61] G. J. Ringer. Letter to the Editor: Quest for bomb literature. Bulletin of the Atomic Scientists, 17(9):393, November 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ringo:1966:LES

[Rin66] Roy Ringo. Letter to the Editor: On strengthening the U.N. Bulletin of the Atomic Scientists, 22(10):30–31, December 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [New67].

Ringo:1975:JSS

[Rin75] Roy Ringo. The justification of science to scientists: A reminder of why the enterprise was launched. Bulletin of the Atomic Scientists, 31(3):29–33, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rob75a].

Ringer:1984:IP

[Rin84a] Dick Ringer. Interdisciplinary programs. Bulletin of the Atomic Scientists, 40(10):21s–23s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ringler:1984:I

[Rin84b] Dick Ringler. Introduction. Bulletin of the Atomic Scientists, 40 (10):5s–6s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ringler:1984:NWT

[Rin84c] Dick Ringler. Nuclear war: a teaching guide. Bulletin of the Atomic Scientists, 40(10):1s–2s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [NR85, You85b].

Ringler:1984:TPN

[Rin84d] Dick Ringler. Toward the practice of nuclear-age education. Bulletin of the Atomic Scientists, 40(10):4s–5s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ripley:1968:ICT

[Rip68] S. Dillon Ripley. International communication in the Twentieth Century. Bulletin of the Atomic Scientists, 24(6):53–55, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riqiang:2015:LMD

[Riq15] Wu Riqiang. Limit missile defense-or expand it? : a Chinese response. Bulletin of the Atomic Scientists, 71(2):9–12, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riskas:1975:LSO

[Ris75] Chris Riskas. Letters: Significant omission. Bulletin of the Atomic Scientists, 31(8):4, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom75a].

Ritchie:1948:SRS

[Rit48] A. D. Ritchie. The special responsibility of scientists. Bulletin of the Atomic Scientists, 4(10):305–306, October 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a, Ash48, Ler48, Ano48v].

Ritterbush:1967:BBA

[Rit67] Philip C. Ritterbush. Books: The Art of Conjecture, by Bertrand de Jouvenel, translation by Nikita Lary. Bulletin of the Atomic Scientists, 23(9):33–34, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rivkin:1962:HW

[Riv62] Steven R. Rivkin. The hobbled weapon. Bulletin of the Atomic Scientists, 18(5):8–13, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Riveles:1996:TSS

[Riv96] Stanley A. Riveles. The treaty is SAFE, says a top U.S. negotiator. Bulletin of the Atomic Scientists, 52(1):25–27, January/ February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reinhardt:1955:TSA

[RK55] George C. Reinhardt and William R. Kintner. The tactical side of atomic warfare. Bulletin of the Atomic Scientists, 11(2):53– 58, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1987:BRB

[RK87a] Thomas Risse-Kappen. Book review: Die Zukunft Europas: Probleme der Friedensgestaltung [(German) The future of Europe: Problems of peace-building], by Dieter Senghaas. Bulletin of the Atomic Scientists, 43(4):50–52, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1987:SWC

[RK87b] Thomas Risse-Kappen. Star Wars controversy in West Germany. Bulletin of the Atomic Scientists, 43(6):50–52, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1988:OGC

[RK88] Thomas Risse-Kappen. Odd German consensus against new missiles. Bulletin of the Atomic Scientists, 44(4):14–17, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1989:BRB

[RK89a] Thomas Risse-Kappen. Book reviews: Pax Atomica: The Nuclear Defense Debate in West Germany during the Adenauer Era, by Mark Cioc; The Silent Partner: West Germany and Arms Control, edited by Barry M. Blechman and Cathleen Fischer. Bulletin of the Atomic Scientists, 45(9):47–48, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1989:WNS

[RK89b] Thomas Risse-Kappen. Will NATO settle for Kohl cuts? Bulletin of the Atomic Scientists, 45(5):9–12, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1990:PNE

[RK90] Thomas Risse-Kappen. Predicting the new Europe. Bulletin of the Atomic Scientists, 46(8):25–29, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1991:NWO

[RK91] Thomas Risse-Kappen. New world order: From Europe, a ray of hope. Bulletin of the Atomic Scientists, 47(5):26–27, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Risse-Kappen:1993:BRC

[RK93] Thomas Risse-Kappen. Book review: Can’t stop seeing red: The Future of Germany and the Atlantic Alliance, by Constantine C. Menges. Bulletin of the Atomic Scientists, 49(1):48–49, January/ February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ramana:2019:CRN

[RK19] M. V. Ramana and Mariia Kurando. Cyberattacks on Russia — the nation with the most nuclear weapons — pose a global threat. Bulletin of the Atomic Scientists, 75(1):44–50, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1556001.

Rogers:2022:LVS

[RKK22] Jessica Rogers, Matt Korda, and Hans M. Kristensen. The long view: Strategic arms control after the New START Treaty. Bulletin of the Atomic Scientists, 78(6):347–368, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2133287.

Reinhardt:1955:CCT

[RKR55] George C. Reinhardt, William R. Kintner, and Eugene I. Rabinowitch. Current comment: The tactical side of atomic warfare. Bulletin of the Atomic Scientists, 11(5):191–192, May 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1969:MM

[RL69] E. Rabinowitch and R. S. Lewis, editors. Man on the Moon. Basic Books, New York, NY, USA, 1969. ???? pp. LCCN ????

Reppy:1976:PPR

[RL76] Judith Reppy and F. A. Long. The Pentagon’s program of R&D: the need for reform. Bulletin of the Atomic Scientists, 32(1):30– 36, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reppy:1978:UMR

[RL78a] Judith Reppy and F. A. Long. U.S. military R&D: a set of questions. Bulletin of the Atomic Scientists, 34(5):34–41, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rudolph:1978:EOC

[RL78b] Lloyd I. Rudolph and Charles S. Lenth. Energy options: changing views from India. Bulletin of the Atomic Scientists, 34(6):6–9, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Red78a].

Rauf:1989:SCB

[RL89] Tariq Rauf and John Lamb. Should Canada bring the boys home? Bulletin of the Atomic Scientists, 45(7):36–38, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sut89].

Rankin-Lowe:1996:LOC

[RL96] Jeff Rankin-Lowe. Letter: Oh Canada! Bulletin of the Atomic Scientists, 52(3):60, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano96k].

Rosner:2014:WAS

[RL14] Robert Rosner and Rebecca Lordan. Why America should move toward dry cask consolidated interim storage of used nuclear fuel. Bulletin of the Atomic Scientists, 70(6):48–62, November 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosner:2011:MPD

[RLG11] Robert Rosner, Rebecca Lordan, and Stephen Goldberg. Moving to passive designs. Bulletin of the Atomic Scientists, 67(4): 23–29, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/4/23.

Rothstein:1998:LLL

[RM98] Linda Rothstein and Brendan Mathews. In a long line of lemons …. Bulletin of the Atomic Scientists, 54(3):31, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ramana:2016:SSN

[RM16] M. V. Ramana and Zia Mian. Scrambling to sell a nuclear Middle East. Bulletin of the Atomic Scientists, 72(1):39–43, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reading:1965:LEG

[RMK65] O. S. Reading, Herbert Malamud, and Don E. Kash. Letter to the Editor: Good science, good politics. Bulletin of the Atomic Scientists, 21(7):25–27, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kas65].

Reistad:2003:RKC

[RMS03] Ole Reistad, Morten Bremer Maerli, and Johnny Skorve. Russia: Kola cleanup coming? Bulletin of the Atomic Scientists, 59(2):12–16, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/2/12.

Reines:1954:WAY

[RMW54] Frederick Reines, Lewis Mumford, and Katharine Way. World assize: Yes or no? Bulletin of the Atomic Scientists, 10(5):159– 161, 168, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roberts:1949:LEF

[Rob49] Arthur Roberts. Letter to the Editor: FAS proposes U.N. distribution of isotopes. Bulletin of the Atomic Scientists, 5(3):79, 95, March 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robbins:1956:BBB

[Rob56] Wallace W. Robbins. Books: The Burning Glass, a play by Charles Morgan. Bulletin of the Atomic Scientists, 12(2):60–61, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robinson:1962:BBA

[Rob62] John J. Robinson. Books: Children of the Ashes, by Robert Jungk. Bulletin of the Atomic Scientists, 18(2):32–34, February 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [JR62].

Robinson:1966:OA

[Rob66] Joan Robinson. The organization of agriculture. Bulletin of the Atomic Scientists, 22(6):28–32, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roberts:1969:BBC

[Rob69] Arthur Roberts. Books: Environment and Change: The Next Fifty Years, by William R. Ewald, Jr. Bulletin of the Atomic Scientists, 25(5):54–55, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roberts:1975:LTM

[Rob75a] Peter Roberts. Letters: Two misconceptions. Bulletin of the Atomic Scientists, 31(8):2–3, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rin75].

Robinson:1975:SCC

[Rob75b] J. P. Perry Robinson. The special case of chemical and biological weapons: Without a breakthrough in disarmament talks, they could be assimilated into world arsenals. Bulletin of the Atomic Scientists, 31(5):17–23, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roberts:1977:LHF

[Rob77] Peter Roberts. Letter: The human factor. Bulletin of the Atomic Scientists, 33(2):64, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei76a].

Roberts:1978:LWT

[Rob78a] Peter Roberts. Letter: A worthless tale. Bulletin of the Atomic Scientists, 34(1):2, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hig77].

Robinson:1978:BHB

[Rob78b] J. P. Perry Robinson. From the broadsword to the hydrogen bomb. Bulletin of the Atomic Scientists, 34(3):44–45, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robinson:1978:NBC

[Rob78c] J. P. Perry Robinson. Neutron bomb and conventional weapons of mass destruction. Bulletin of the Atomic Scientists, 34(3): 42–45, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robinowitz:1988:LSS

[Rob88] Mark Robinowitz. Letter: Spy satellite bill. Bulletin of the Atomic Scientists, 44(9):45–46, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Str88].

Robock:1989:NMC

[Rob89] Alan Robock. New models confirm nuclear winter. Bulletin of the Atomic Scientists, 45(7):32, 34–35, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robinson:1990:LBF

[Rob90a] Brian Robinson. Letter: Bulletin flunked geography …. Bulletin of the Atomic Scientists, 46(7):47, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robinwitz:1990:LDU

[Rob90b] Mark Robinwitz. Letter: Defend U.S. shores and the environment. Bulletin of the Atomic Scientists, 46(4):54, May 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robinowitz:1991:LGP

[Rob91a] Mark Robinowitz. Letter: Green Pentagon sheer Nunn-sense. Bulletin of the Atomic Scientists, 47(2):47, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa90x].

Robock:1991:BRI

[Rob91b] Alan Robock. Book review: The imparsible dream? A Path Where No Man Thought: Nuclear Winter and the End of the Arms Race, by Carl Sagan and Richard Turco. Bulletin of the Atomic Scientists, 47(2):43–44, March 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robock:2008:RWG

[Rob08] Alan Robock. 20 reasons why geoengineering may be a bad idea. Bulletin of the Atomic Scientists, 64(2):14–18, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/2/14.

Roberts:2018:WSB

[Rob18] Thomas G. Roberts. Why a space-based missile interceptor system is not viable. Bulletin of the Atomic Scientists, 74(4): 238–242, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1486597.

Roberts:2020:TJN

[Rob20] Brad Roberts. It’s time to jettison Nuclear Posture Reviews. Bulletin of the Atomic Scientists, 76(1):31–36, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1701282.

Rochlin:1976:RSN

[Roc76] Gene Rochlin. Rebuttal: Secret nuclear weapons programs could be difficult to detect. Bulletin of the Atomic Scientists, 32(5):63, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gla76a].

Rochlin:1978:LWV

[Roc78] Gene I. Rochlin. Lessons from West Valley. Bulletin of the Atomic Scientists, 34(1):24, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rochlin:1981:BRS

[Roc81] Gene Rochlin. Book reviews: Strategic doctrines: Force without War: U.S. Armed Forces as a Political Instrument, by Barry Blechman and Stephen S. Kaplan; The Dangers of Nuclear War, edited by Franklyn Griffiths and John C. Polanyi; The Nuclear Question: The United States and Nuclear Weapons, 1946–1976, by Michael Mandelbaum; Strategic Though in the Nuclear Age, edited by Laurence Martin; Limited War Revisited, by Robert E. Osgood. Bulletin of the Atomic Scientists, 37(2):48–54, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rochlin:1983:BRB

[Roc83a] Gene I. Rochlin. Book reviews: The Evolution of Nuclear Strategy, by Lawrence Freedman; Strategic Studies and Public Policy: The American Experience, by Colin S. Gray. Bulletin of the Atomic Scientists, 39(10):42–44, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum correcting publisher of second book [Ano84h].

Rochlin:1983:CCT

[Roc83b] Gene I. Rochlin. Commentary: On ‘The connection is tenuous’. Bulletin of the Atomic Scientists, 39(7):64, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spi83].

Rochlin:1985:BRB

[Roc85] Gene I. Rochlin. Book review: State of the World: 1984, by Lester R. Brown, William Chandler, Christopher Flavin, Sandra Postel, Linda Starke, and Edward Wolfe; State of the World: 1985, by Lester R. Brown, William Chandler, Christopher Flavin, Sandra Postel, Linda Starke, Edward Wolfe, and Cynthia Pollock. Bulletin of the Atomic Scientists, 41(6):45–46, June/July 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roche:1997:LTP

[Roc97a] Douglas Roche. Letter: A time for positive pressure. Bulletin of the Atomic Scientists, 53(6):62–63, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kri97].

Rockwell:1997:LEE

[Roc97b] Theodore Rockwell. Letter: Exposure explained. Bulletin of the Atomic Scientists, 53(6):64, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano97l].

Rockwood:2018:HIV

[Roc18] Laura Rockwood. How the IAEA verifies if a country’s nuclear program is peaceful or not: The legal basis. Bulletin of the Atomic Scientists, 74(5):317–325, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1507789.

Rodgers:1961:NGI

[Rod61] Raymond Spencer Rodgers. Non-governmental IAEA representatives. Bulletin of the Atomic Scientists, 17(1):24, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rodberg:1967:ASA

[Rod67] Leonard S. Rodberg. ABM — some arms control issues. Bulletin of the Atomic Scientists, 23(6):16–20, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rodberg:1969:LST

[Rod69] Leonard S. Rodberg. Limiting strategic technology: The need for national self-restraint. Bulletin of the Atomic Scientists, 25 (9):36–38, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rodberg:1970:BBW

[Rod70] Leonard S. Rodberg. Books: Report from Wasteland: America’s Military–Industrial Complex, by Senator William Proxmire. Bulletin of the Atomic Scientists, 26(7):46–47, September 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rodberg:1985:BRB

[Rod85] Leonard S. Rodberg. Book reviews: The Wizards of Armageddon, by Fred Kaplan; Counsels of War. Bulletin of the Atomic Scientists, 41(9):44–46, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roederer:1991:LTS

[Roe91] Juan G. Roederer. Let a thousand Sakharovs bloom. Bulletin of the Atomic Scientists, 47(6):18–21, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roff:1997:GCP

[Rof97] Sue Rabbitt Roff. The ghost of Christmas past. Bulletin of the Atomic Scientists, 53(5):52–56, September/October 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano97q].

Roff:2016:BRA

[Rof16] Heather Roff. To ban or regulate autonomous weapons: A US response. Bulletin of the Atomic Scientists, 72(2):122–124, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roff:2019:FPA

[Rof19] Heather M. Roff. The frame problem: The AI “arms race” isn’t one. Bulletin of the Atomic Scientists, 75(3):95–98, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1604836.

Rogers:1971:UNP

[Rog71] Franklyn C. Rogers. Underground nuclear power plants. Bulletin of the Atomic Scientists, 27(8):38–41, 51, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rogers:1988:NC

[Rog88a] Paul Rogers. The nuclear connection. Bulletin of the Atomic Scientists, 44(7):20–23, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bru88].

Rogoway:1988:TSS

[Rog88b] Raymond Rogoway. Teaching science and social responsibility. Bulletin of the Atomic Scientists, 44(6):19, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rollefson:1957:WMM

[Rol57] R. Rollefson. Why so many missiles? Bulletin of the Atomic Scientists, 13(8):295–301, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roling:1980:CSS

[Röl80] Bert V. A. Röling. Commentary: The sin of silence. Bulletin of the Atomic Scientists, 36(9):10–13, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roling:1981:JV

[Röl81] Bert V. A. Röling. A judge’s view. Bulletin of the Atomic Scientists, 37(8):52–53, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Romancheck:1977:LNS

[Rom77] Robert Romancheck. Letter: No solar subsidies. Bulletin of the Atomic Scientists, 33(1):2, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hir76, Car77].

Romm:1982:BRB

[Rom82] Joseph J. Romm. Book review: Life After Nuclear War, by Arthur Katz. Bulletin of the Atomic Scientists, 38(6):48–49, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Romm:1983:WTP

[Rom83] Joseph J. Romm. The weapons tutorial: [part 6:] the gaps of wrath. Bulletin of the Atomic Scientists, 39(8):28–32, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Parts 1 [Mor83c], 2 [Tsi83i], 3 [Mar83b], 4 [Hol83c], and 5 [Sar83].

Romm:1989:BRB

[Rom89] Joseph J. Romm. Book review: Infinite in All Directions, by Freeman Dyson. Bulletin of the Atomic Scientists, 45(8):35–36, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Romm:1990:BRM

[Rom90] Joseph J. Romm. Book review: Myths and misses Science à la Mode: Physical Fashions and Fictions, by Tony Rothman. Bulletin of the Atomic Scientists, 46(5):39–40, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Romm:1991:NNR

[Rom91] Joseph Romm. Needed — a no-regrets energy policy. Bulletin of the Atomic Scientists, 47(6):31–36, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Romm:1992:LWW

[Rom92] Joseph J. Romm. Laid waste by weapons lust. Bulletin of the Atomic Scientists, 48(8):15–23, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ropeik:2016:DR

[Rop16] David Ropeik. The dangers of radiophobia. Bulletin of the Atomic Scientists, 72(5):311–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rostand:1952:BBO

[Ros52] Jean Rostand. Biology and the burden of our times. Bulletin of the Atomic Scientists, 8(6):176–178, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Olt52].

Rosenfeld:1959:TBT

[Ros59] Arthur H. Rosenfeld. Test ban: Two views: What about the undetectable tests? Bulletin of the Atomic Scientists, 15(3):98, 103–108, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosebury:1960:SHC

[Ros60] Theodor Rosebury. Some historical considerations. Bulletin of the Atomic Scientists, 16(6):227–236, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rostlund:1961:TT

[Ros61] Erhard Rostlund. Taming trees. Bulletin of the Atomic Scientists, 17(8):326–330, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosa:1962:LEB

[Ros62] Nicholas Rosa. Letter to the Editor: Buying time. Bulletin of the Atomic Scientists, 18(9):32, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys62].

Roshwald:1963:IF

[Ros63] Mordecai Roshwald. An international force. Bulletin of the Atomic Scientists, 19(4):24–27, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenthal:1967:BBR

[Ros67] Gerson M. Rosenthal, Jr. Books: The Reforming of General Education, by Daniel Bell. Bulletin of the Atomic Scientists, 23 (2):23–24, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenbloom:1968:BBC

[Ros68a] Richard S. Rosenbloom. Books: Technology and Change, by Donald Schon. Bulletin of the Atomic Scientists, 24(7):27–28, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenstock:1968:CCW

[Ros68b] Herbert B. Rosenstock. Current comments: Who should support scholarly research? Bulletin of the Atomic Scientists, 24(1):60– 62, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roszak:1969:CDA

[Ros69] Theodore Roszak. Communications: Dissenting academician. Bulletin of the Atomic Scientists, 25(8):46–47, October 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hym68, Hym69b].

Rose:1975:SVA

[Ros75] William M. Rose. Submarine vs. anti-submarine: The billiondollar merry-go-round: The Navy pushes for Trident, but there’s a better way to keep its sea-based deterrent invulnerable. Bulletin of the Atomic Scientists, 31(4):27–30, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano75d].

Rosa:1978:PEP

[Ros78a] Eugene Rosa. The public and the energy problem. Bulletin of the Atomic Scientists, 34(4):5–7, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ross:1978:CCP

[Ros78b] Douglas Ross. The Concorde compromise: the politics of decision making. Bulletin of the Atomic Scientists, 34(3):46–53, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sto78b].

Rosenberg:1981:CDR

[Ros81] Milton J. Rosenberg. Commentary: The decline and rise of the Cold War consensus. Bulletin of the Atomic Scientists, 37(3): 7–9, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenberg:1982:USN

[Ros82] David Alan Rosenberg. United States nuclear stockpile, 1945 to 1950. Bulletin of the Atomic Scientists, 38(5):25–30, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rose:1985:NP

[Ros85] David J. Rose. Nuclear power. Bulletin of the Atomic Scientists, 41(7):76–78, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [PR85].

Rosenberg:1987:UBW

[Ros87a] Barbara Hatch Rosenberg. Updating the biological weapons ban. Bulletin of the Atomic Scientists, 43(1):40–43, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenberg:1987:UNS

[Ros87b] David Alan Rosenberg. U.S. nuclear strategy: theory vs. practice. Bulletin of the Atomic Scientists, 43(2):20–26, March 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenberg:1988:WRM

[Ros88] Raoul Rosenberg. Washington report: The many phases of SDI. Bulletin of the Atomic Scientists, 44(4):3–4, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosen:1989:PPH

[Ros89a] Jay Rosen. Phantom public haunts nuclear age. Bulletin of the Atomic Scientists, 45(5):16–19, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [LKWR89].

Ross:1989:RMB

[Ros89b] Michael Ross. Report: Military budget: Trident II misfires in Congress. Bulletin of the Atomic Scientists, 45(10):11–12, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosen:1990:BRE

[Ros90a] Jay Rosen. Book review: Exercising the public veto: The Rockets’ Red Glare: When America Goes to War — the Presidents and the People, by Richard J. Barnet. Bulletin of the Atomic Scientists, 46(6):42–43, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ross:1990:DIY

[Ros90b] Andrew L. Ross. Do-it-yourself weaponry. Bulletin of the Atomic Scientists, 46(4):20–22, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenberg:1994:NLW

[Ros94] Barbara Hatch Rosenberg. “Non-lethal” weapons may violate treaties. Bulletin of the Atomic Scientists, 50(5):44–45, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano94s].

Rossin:1998:LUL

[Ros98] A. David Rossin. Letter: U.S. losing ground. Bulletin of the Atomic Scientists, 54(1):67–68, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenberg:2006:WWT

[Ros06a] Barbara Hatch Rosenberg. Watching while they work. Bulletin of the Atomic Scientists, 62(6):64, November/December 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/6/64.

Rosenberg:2006:FF

[Ros06b] Emily Rosenberg. Friendly fire. Bulletin of the Atomic Scientists, 62(4):57–59, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/4/57.

Rosner:2008:MNE

[Ros08] Robert Rosner. Making nuclear energy work. Bulletin of the Atomic Scientists, 64(1):28–33, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/1/28.

Rotblat:1955:HUB

[Rot55] J. Rotblat. The hydrogen–uranium bomb. Bulletin of the Atomic Scientists, 11(5):171–172, 177, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1958:SNW

[Rot58a] Joseph Rotblat. Supplement: Nuclear weapons in war: Nuclear weapons in war. Bulletin of the Atomic Scientists, 14(1):57–58, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotkin:1958:RCS

[Rot58b] Israel Rotkin. Reader comment: Science education for all. Bulletin of the Atomic Scientists, 14(9):400, November 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothschild:1961:RCC

[Rot61] Jacquard H. Rothschild. Reflections and comments: Can peace be guaranteed? Bulletin of the Atomic Scientists, 17(9):387– 389, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bec62].

Rotblat:1962:BBT

[Rot62] Joseph Rotblat. Books: Current Trends in Scientific Research, by Pierre Auger. Bulletin of the Atomic Scientists, 18(3):40–41, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roth:1972:CSM

[Rot72] Dana L. Roth. Communications: Science in Mainland China. Bulletin of the Atomic Scientists, 28(7):3, September 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Esp72b].

Rotblat:1977:CWG

[Rot77] Joseph Rotblat. Controlling weapons-grade fissile material: at present growth rates, the world’s commercial nuclear reactors by 1990 will be producing enough plutonium for 300,000 atomic bombs; tightening of the safeguards should no longer be postponed. Bulletin of the Atomic Scientists, 33(6):37–43, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1978:NEI

[Rot78a] Joseph Rotblat. Nuclear energy and independence. Bulletin of the Atomic Scientists, 34(6):19–20, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1978:RRW

[Rot78b] Joseph Rotblat. The risks for radiation workers. Bulletin of the Atomic Scientists, 34(7):41–46, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CMR79, Coh79].

Rotblat:1979:BRB

[Rot79a] Joseph Rotblat. Book reviews: Review of the Department of Energy’s Controversial Termination of a Research Contract; Radiation Standards and Public Safety. Bulletin of the Atomic Scientists, 35(8):39–41, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See apology and comment [AR80].

Rotblat:1979:EPW

[Rot79b] Joseph Rotblat. Einstein the pacifist warrior. Bulletin of the Atomic Scientists, 35(3):21–26, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Rotblat:1981:HLL

[Rot81a] Joseph Rotblat. Hazards of low-level radiation — less agreement, more confusion. Bulletin of the Atomic Scientists, 37(6):31–36, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gro82].

Rotblat:1981:REM

[Rot81b] Joseph Rotblat. The Russell–Einstein Manifesto 25 years later: The threat today. Bulletin of the Atomic Scientists, 37(1):33– 36, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothschild:1983:IYL

[Rot83] Edwin A. Rothschild. If you love these films. Bulletin of the Atomic Scientists, 39(6):39, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Roth:1984:CP

[Rot84] Eli B. Roth. Commentary: Proliferation. Bulletin of the Atomic Scientists, 40(7):54, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WL84].

Rotblat:1985:LBP

[Rot85] Joseph Rotblat. Leaving the bomb project. Bulletin of the Atomic Scientists, 41(7):16–19, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1988:ASB

[Rot88] Joseph Rotblat. After the summit: British fret about “vulnerability”. Bulletin of the Atomic Scientists, 44(2):20–22, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1989:BRC

[Rot89a] Jandos Rothstein. Book review: Cockroaches and pyramids: Quonset Huts on the River Styx: the Bomb Shelter Design Book. Bulletin of the Atomic Scientists, 45(10):46–47, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1989:BVA

[Rot89b] Linda Rothstein. Biotech ventures in academia — have they paid off? Bulletin of the Atomic Scientists, 45(3):34, April 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1991:BWW

[Rot91a] Linda Rothstein. Bulletin: War of words over women warriors. Bulletin of the Atomic Scientists, 47(10):6–7, December 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1991:RAR

[Rot91b] Linda Rothstein. Report: Arms reductions: Cut, but carefully. Bulletin of the Atomic Scientists, 47(9):11, 46, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1992:CV

[Rot92a] Joseph Rotblat. Citizen verification. Bulletin of the Atomic Scientists, 48(4):18–20, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:SNC

[Rot92b] Linda Rothstein. 7,000,000 secrets; no common sense. Bulletin of the Atomic Scientists, 48(5):5–6, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BBW

[Rot92c] Linda Rothstein. Big brother is watching — TV. Bulletin of the Atomic Scientists, 48(5):5, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BBB

[Rot92d] Linda Rothstein. Bulletin: Buzz buzz. Bulletin of the Atomic Scientists, 48(3):5–6, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BDM

[Rot92e] Linda Rothstein. Bulletin: Drumming up (military) business? Bulletin of the Atomic Scientists, 48(6):5–6, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BFF

[Rot92f] Linda Rothstein. Bulletin: Farmers fear booming business. Bulletin of the Atomic Scientists, 48(8):4, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BGO

[Rot92g] Linda Rothstein. Bulletin: George Orwell would have understood. Bulletin of the Atomic Scientists, 48(4):6, May 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BHT

[Rot92h] Linda Rothstein. Bulletin: Haven’t they seen a single science fiction film? Bulletin of the Atomic Scientists, 48(7):7, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BHY

[Rot92i] Linda Rothstein. Bulletin: Heads you pay; tails, we don’t. Bulletin of the Atomic Scientists, 48(9):5–6, November 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BBP

[Rot92j] Linda Rothstein. Bulletin: It’s a bird — it’s a plane — it’s the black budget. Bulletin of the Atomic Scientists, 48(6):4–5, July/ August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BME

[Rot92k] Linda Rothstein. Bulletin: Military exercises; musical chairs. Bulletin of the Atomic Scientists, 48(8):3–4, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BMR

[Rot92l] Linda Rothstein. Bulletin: More roadblocks to chemical treaty? Bulletin of the Atomic Scientists, 48(5):7, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BND

[Rot92m] Linda Rothstein. Bulletin: Nevadans dump dump ads. Bulletin of the Atomic Scientists, 48(4):3–4, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BNM

[Rot92n] Linda Rothstein. Bulletin: No matter what, says NAS, atomic vets are O.K. Bulletin of the Atomic Scientists, 48(10):3–4, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BNP

[Rot92o] Linda Rothstein. Bulletin: No party for Star Wars. Bulletin of the Atomic Scientists, 48(5):3–4, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BNY

[Rot92p] Linda Rothstein. Bulletin: Now you see it, now you don’t. Bulletin of the Atomic Scientists, 48(3):6, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BRR

[Rot92q] Linda Rothstein. Bulletin: Russian roulette. Bulletin of the Atomic Scientists, 48(8):5, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fin92].

Rothstein:1992:BWT

[Rot92r] Linda Rothstein. Bulletin: So why not tell us? Bulletin of the Atomic Scientists, 48(7):6, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BWL

[Rot92s] Linda Rothstein. Bulletin: The weak link. Bulletin of the Atomic Scientists, 48(6):7, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BVP

[Rot92t] Linda Rothstein. Bulletin: “Vampire phenomenon”. Bulletin of the Atomic Scientists, 48(7):6–7, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:BVH

[Rot92u] Linda Rothstein. Bulletin: Vitrification — how booming a business? Bulletin of the Atomic Scientists, 48(9):6–7, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1992:ISS

[Rot92v] Linda Rothstein. International spoil sports. Bulletin of the Atomic Scientists, 48(10):5, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BRU

[Rot93a] Linda Rothstein. Book review: U and Pu: World Inventory of Plutonium and Highly Enriched Uranium, 1992, by David Albright, Frans Berkhout, and William Walker. Bulletin of the Atomic Scientists, 49(6):52–53, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BAS

[Rot93b] Linda Rothstein. Bulletin: “Amerika for sale”. Bulletin of the Atomic Scientists, 49(2):5, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BAA

[Rot93c] Linda Rothstein. Bulletin: Arms as alms. Bulletin of the Atomic Scientists, 49(8):6, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BBT

[Rot93d] Linda Rothstein. Bulletin: Budget testing. Bulletin of the Atomic Scientists, 49(8):5, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BWW

[Rot93e] Linda Rothstein. Bulletin: But who will run the new generation? Bulletin of the Atomic Scientists, 49(1):5, January/ February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BDW

[Rot93f] Linda Rothstein. Bulletin: Defense work means jobs — for Mexicans. Bulletin of the Atomic Scientists, 49(9):5, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BDD

[Rot93g] Linda Rothstein. Bulletin: Drip, drip, drip. Bulletin of the Atomic Scientists, 49(8):5, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BFM

[Rot93h] Linda Rothstein. Bulletin: FEMA, the man-man disaster. Bulletin of the Atomic Scientists, 49(6):7–8, July/August 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BBE

[Rot93i] Linda Rothstein. Bulletin: For British eyes only? Bulletin of the Atomic Scientists, 49(3):6, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BHM

[Rot93j] Linda Rothstein. Bulletin: Hail to the Military Plenipotentiary. Bulletin of the Atomic Scientists, 49(2):3–4, March 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BHS

[Rot93k] Linda Rothstein. Bulletin: Holy smokes, Batman! Bulletin of the Atomic Scientists, 49(2):5, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BKI

[Rot93l] Linda Rothstein. Bulletin: Keeping “it” down on the farm. Bulletin of the Atomic Scientists, 49(4):8, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BMH

[Rot93m] Linda Rothstein. Bulletin: Making hay bashing gays. Bulletin of the Atomic Scientists, 49(4):6–7, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BMT

[Rot93n] Linda Rothstein. Bulletin: Moving targets. Bulletin of the Atomic Scientists, 49(7):7, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BPN

[Rot93o] Linda Rothstein. Bulletin: Plugging the nuclear pipeline. Bulletin of the Atomic Scientists, 49(9):4–5, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BPL

[Rot93p] Linda Rothstein. Bulletin: Preserving losers’ history. Bulletin of the Atomic Scientists, 49(5):5–6, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BRS

[Rot93q] Linda Rothstein. Bulletin: Run, SPOT, run. Bulletin of the Atomic Scientists, 49(7):4–5, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BTW

[Rot93r] Linda Rothstein. Bulletin: Tearing wings off flies. Bulletin of the Atomic Scientists, 49(2):6, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BWS

[Rot93s] Linda Rothstein. Bulletin: The wages of sin? About $4.1 million. Bulletin of the Atomic Scientists, 49(1):5, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:BWB

[Rot93t] Linda Rothstein. Bulletin: What a blast. Bulletin of the Atomic Scientists, 49(10):5, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano94r].

Rothstein:1993:BYC

[Rot93u] Linda Rothstein. Bulletin: Y’all come down to the missile shoot. Bulletin of the Atomic Scientists, 49(10):4, December 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:KB

[Rot93v] Linda Rothstein. Kapitsa on Beria. Bulletin of the Atomic Scientists, 49(4):27, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1993:RRH

[Rot93w] Linda Rothstein. Report: Radiation and health: PSR pinpoints problems. Bulletin of the Atomic Scientists, 49(1):10, 50–51, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1994:LTS

[Rot94a] Joseph Rotblat. Letter: Ten steps to peace. Bulletin of the Atomic Scientists, 50(3):60–61, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TM94c].

Rothstein:1994:BCT

[Rot94b] Linda Rothstein. Bulletin: Crawling toward CRADAs. Bulletin of the Atomic Scientists, 50(5):9, September/October 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BDB

[Rot94c] Linda Rothstein. Bulletin: Danger — bureaucrats at work. Bulletin of the Atomic Scientists, 50(6):8, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BFN

[Rot94d] Linda Rothstein. Bulletin: French nuclear power loses its punch. Bulletin of the Atomic Scientists, 50(4):8–9, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bes94].

Rothstein:1994:BGG

[Rot94e] Linda Rothstein. Bulletin: Greenpeace gets the goods. Bulletin of the Atomic Scientists, 50(6):6–7, November/December 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BOW

[Rot94f] Linda Rothstein. Bulletin: Oh, what a tangled web we weave. Bulletin of the Atomic Scientists, 50(4):6–7, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BLF

[Rot94g] Linda Rothstein. Bulletin: The land of the free … bullet. Bulletin of the Atomic Scientists, 50(1):8, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BSK

[Rot94h] Linda Rothstein. Bulletin: The ‘soft kill’ solution. Bulletin of the Atomic Scientists, 50(2):4–6, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BTC

[Rot94i] Linda Rothstein. Bulletin: They could have just asked Mom. Bulletin of the Atomic Scientists, 50(5):8–9, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:BWS

[Rot94j] Linda Rothstein. Bulletin: Whoops, the series. Bulletin of the Atomic Scientists, 50(1):9, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1994:SWR

[Rot94k] Linda Rothstein. Star Wars redux. Bulletin of the Atomic Scientists, 50(3):9–10, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:TP

[Rot95a] Linda Rothstein. “177 tanks, 177 problems”. Bulletin of the Atomic Scientists, 51(3):39, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BLT

[Rot95b] Linda Rothstein. Bulletin: A little token of appreciation. Bulletin of the Atomic Scientists, 51(5):7–8, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BMD

[Rot95c] Linda Rothstein. Bulletin: A matter of definition. Bulletin of the Atomic Scientists, 51(4):8–9, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BCP

[Rot95d] Linda Rothstein. Bulletin: Chemists, physicists duke it out. Bulletin of the Atomic Scientists, 51(4):7, July/August 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BOT

[Rot95e] Linda Rothstein. Bulletin: Operation tricky nickname. Bulletin of the Atomic Scientists, 51(6):7, November/December 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BRR

[Rot95f] Linda Rothstein. Bulletin: Reagan’s revenge? Bulletin of the Atomic Scientists, 51(6):8–9, November/December 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BSR

[Rot95g] Linda Rothstein. Bulletin: “Secrets” off the rack. Bulletin of the Atomic Scientists, 51(2):6–7, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BST

[Rot95h] Linda Rothstein. Bulletin: Sitting tall. Bulletin of the Atomic Scientists, 51(1):8–9, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:BTW

[Rot95i] Linda Rothstein. Bulletin: The transfermium wars. Bulletin of the Atomic Scientists, 51(1):5–6, January/February 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:EEF

[Rot95j] Linda Rothstein. The egregious eight: a few programs to keep an eye on. Bulletin of the Atomic Scientists, 51(5):44–45, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:NCA

[Rot95k] Linda Rothstein. Nothing clean about “cleanup”. Bulletin of the Atomic Scientists, 51(3):34–35, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Photos by Robert Del Tredici.

Rothstein:1995:PYH

[Rot95l] Linda Rothstein. Perspective: Yes, haste made waste. Bulletin of the Atomic Scientists, 51(3):14–15, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1995:DR

[Rot95m] Linda Rothstein. That distant rumble. Bulletin of the Atomic Scientists, 51(3):7, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1996:RYH

[Rot96a] Joseph Rotblat. Remember your humanity. Bulletin of the Atomic Scientists, 52(2):26–28, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BRB

[Rot96b] Linda Rothstein. Book review: $1 billion a day: When the Pentagon Was for Sale, by Andy Pasztor. Bulletin of the Atomic Scientists, 52(4):59–60, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BCY

[Rot96c] Linda Rothstein. Bulletin: Cassini, yes; Topaz, no. Bulletin of the Atomic Scientists, 52(5):9–10, September/October 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BEE

[Rot96d] Linda Rothstein. Bulletin: Excuses, excuses. Bulletin of the Atomic Scientists, 52(6):10, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BGN

[Rot96e] Linda Rothstein. Bulletin: Good news, bad news. Bulletin of the Atomic Scientists, 52(6):8–9, November/December 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BHA

[Rot96f] Linda Rothstein. Bulletin: Hey, army, welcome to our world. Bulletin of the Atomic Scientists, 52(4):5–6, July/August 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BJC

[Rot96g] Linda Rothstein. Bulletin: Jumping to conclusions. Bulletin of the Atomic Scientists, 52(4):10, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BLM

[Rot96h] Linda Rothstein. Bulletin: Looking into Mitre’s corners. Bulletin of the Atomic Scientists, 52(3):7–8, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BPT

[Rot96i] Linda Rothstein. Bulletin: Physicists test threshold — of tolerance. Bulletin of the Atomic Scientists, 52(2):5–6, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BPS

[Rot96j] Linda Rothstein. Bulletin: Pu in the sky. Bulletin of the Atomic Scientists, 52(1):6, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BSR

[Rot96k] Linda Rothstein. Bulletin: Searching for reality. Bulletin of the Atomic Scientists, 52(4):9–10, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1996:BFF

[Rot96l] Linda Rothstein. Bulletin: The Flipper flap. Bulletin of the Atomic Scientists, 52(5):8, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rotblat:1997:LLS

[Rot97a] Joseph Rotblat. Letter: Limits on science. Bulletin of the Atomic Scientists, 53(4):60, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BRPa

[Rot97b] Linda Rothstein. Book review: Policing the past: History Wars, edited by Edward T. Linenthal and Tom Engelhardt. Bulletin of the Atomic Scientists, 53(1):63–64, January/February 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BRPb

[Rot97c] Linda Rothstein. Book review: The plane truth: An Exhibit Denied: Lobbying the History of the Enola Gay, by Martin Harwit, Springer-Verlag New York, 1996, 502 pages, $27.50. Bulletin of the Atomic Scientists, 53(2):55–57, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BO

[Rot97d] Linda Rothstein. Bullatomsci.org. Bulletin of the Atomic Scientists, 53(5):8, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BFM

[Rot97e] Linda Rothstein. Bulletin: Fissile materials: Facts & figures. Bulletin of the Atomic Scientists, 53(3):9–10, May/June 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BSR

[Rot97f] Linda Rothstein. Bulletin: For science, right or wrong. Bulletin of the Atomic Scientists, 53(1):8, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BHB

[Rot97g] Linda Rothstein. Bulletin: Happy birthday, Mr. Lodestar. Bulletin of the Atomic Scientists, 53(3):8–9, May/June 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BRW

[Rot97h] Linda Rothstein. Bulletin: Russia’s weapons: From the bizarre to the bazaar. Bulletin of the Atomic Scientists, 53(4):5–7, July/ August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BTT

[Rot97i] Linda Rothstein. Bulletin: Turbulent times for fusion power. Bulletin of the Atomic Scientists, 53(5):5–8, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:BUU

[Rot97j] Linda Rothstein. Bulletin: U.S., Ukraine at cross porpoises. Bulletin of the Atomic Scientists, 53(6):8–9, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1997:GTR

[Rot97k] Linda Rothstein. The great tritium rush. Bulletin of the Atomic Scientists, 53(2):9–10, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sto97].

Rothstein:1998:BRSa

[Rot98a] Linda Rothstein. Book review: Secrets and saucer nuts: Area 51: The Dreamland Chronicles, by David Darlington. Bulletin of the Atomic Scientists, 54(3):64–65, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:BAR

[Rot98b] Linda Rothstein. Bulletin: All in the (Russian) mind? Bulletin of the Atomic Scientists, 54(4):11, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:BIW

[Rot98c] Linda Rothstein. Bulletin: An idea whose time has come — and gone? Bulletin of the Atomic Scientists, 54(4):7–11, July/ August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:BMS

[Rot98d] Linda Rothstein. Bulletin: Is making sense too much to ask? Bulletin of the Atomic Scientists, 54(2):5–6, March/April 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:BRSb

[Rot98e] Linda Rothstein. Bulletin: Reaching for the stars. Bulletin of the Atomic Scientists, 54(5):9, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Far98].

Rothstein:1998:BPW

[Rot98f] Linda Rothstein. Bulletin: “The problem was not hidden”. Bulletin of the Atomic Scientists, 54(5):6–7, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:BVS

[Rot98g] Linda Rothstein. Bulletin: Vanunu sprung from solitary. Bulletin of the Atomic Scientists, 54(3):8, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:LSL

[Rot98h] Linda Rothstein. Letter: Shedding light on the landmine problem. Bulletin of the Atomic Scientists, 54(6):10–11, November/ December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:PNG

[Rot98i] Linda Rothstein. Playing the numbers game. Bulletin of the Atomic Scientists, 54(3):32, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1998:SNO

[Rot98j] Linda Rothstein. She’s number one! she’s number one! Bulletin of the Atomic Scientists, 54(3):10–11, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1999:CBC

[Rot99a] Linda Rothstein. The crystal ball, crack’d. Bulletin of the Atomic Scientists, 55(4):10, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:1999:NSP

[Rot99b] Linda Rothstein. Nuclear secrets: The Perseus Papers. Bulletin of the Atomic Scientists, 55(4):17–19, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/4/17.

Rothstein:1999:SM

[Rot99c] Linda Rothstein. The sound of music. Bulletin of the Atomic Scientists, 55(5):12, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:2000:MTM

[Rot00a] Linda Rothstein. More than marketing. Bulletin of the Atomic Scientists, 56(5):2, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/5/2.

Rothstein:2000:U

[Rot00b] Linda Rothstein. “Unfortunate”. Bulletin of the Atomic Scientists, 56(4):2, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/4/2.

Rothstein:2000:WWC

[Rot00c] Linda Rothstein. Where we came in. Bulletin of the Atomic Scientists, 56(6):2, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/6/2.

Rothstein:2001:AS

[Rot01a] Linda Rothstein. After September 11. Bulletin of the Atomic Scientists, 57(6):44–49, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/44.

Rothstein:2001:CW

[Rot01b] Linda Rothstein. Creative writing. Bulletin of the Atomic Scientists, 57(2):2, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/2/2.

Rothstein:2001:HAR

[Rot01c] Linda Rothstein. Humor amid the rubble. Bulletin of the Atomic Scientists, 57(6):2, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/2.

Rothstein:2001:PC

[Rot01d] Linda Rothstein. Plutonium, the contest. Bulletin of the Atomic Scientists, 57(3):28–29, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/3/28.

Rothstein:2001:Pb

[Rot01e] Linda Rothstein. Powering up. Bulletin of the Atomic Scientists, 57(4):2, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/4/2.

Rothstein:2001:Pa

[Rot01f] Linda Rothstein. Predictions. Bulletin of the Atomic Scientists, 57(1):2, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/1/2.

Rothstein:2001:PP

[Rot01g] Linda Rothstein. The purloined plutonium. Bulletin of the Atomic Scientists, 57(5):2, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/5/2.

Rothstein:2001:TC

[Rot01h] Linda Rothstein. Things change. Bulletin of the Atomic Scientists, 57(3):2, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/57/3/2.

Rothstein:2002:RAA

[Rot02a] Linda Rothstein. Not really alike at all. Bulletin of the Atomic Scientists, 58(3):2, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/3/2.

Rothstein:2002:NI

[Rot02b] Linda Rothstein. Nuclear insecurity. Bulletin of the Atomic Scientists, 58(1):2, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/1/2.

Rothstein:2002:OWD

[Rot02c] Linda Rothstein. One weird dude. Bulletin of the Atomic Scientists, 58(5):67–68, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/5/67.

Rothstein:2002:PS

[Rot02d] Linda Rothstein. Phony stories. Bulletin of the Atomic Scientists, 58(5):2, September/October 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/5/2.

Rothstein:2002:SVS

[Rot02e] Linda Rothstein. Safety v. security. Bulletin of the Atomic Scientists, 58(2):1, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/2/1.

Rothstein:2002:UA

[Rot02f] Linda Rothstein. The unthinkable, again. Bulletin of the Atomic Scientists, 58(4):2, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/4/2.

Rothstein:2002:DA

[Rot02g] Linda Rothstein. With dominion over all. Bulletin of the Atomic Scientists, 58(6):2, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/6/2.

Rothstein:2003:AH

[Rot03a] Linda Rothstein. Ah, hubris. Bulletin of the Atomic Scientists, 59(3):2, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/3/2.

Rothstein:2003:BIG

[Rot03b] Linda Rothstein. Bad ideas, good articles. Bulletin of the Atomic Scientists, 59(6):2, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/6/2.

Rothstein:2003:CPH

[Rot03c] Linda Rothstein. A call for a public hearing. Bulletin of the Atomic Scientists, 59(5):2, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/5/2.

Rothstein:2003:LF

[Rot03d] Linda Rothstein. Loyal to a fault. Bulletin of the Atomic Scientists, 59(4):2, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/4/2.

Rothstein:2003:TT

[Rot03e] Linda Rothstein. Transferring terror. Bulletin of the Atomic Scientists, 59(2):2, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/2/2.

Rothstein:2003:WS

[Rot03f] Linda Rothstein. Why not Switzerland? Bulletin of the Atomic Scientists, 59(1):2, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/2.

Rothstein:2004:PP

[Rot04a] Jandos Rothstein. Pretty poison. Bulletin of the Atomic Scientists, 60(4):38–39, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/4/38.

Rothstein:2004:BB

[Rot04b] Linda Rothstein. Blowing bubbles. Bulletin of the Atomic Scientists, 60(1):2, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/60/1/2.

Rothstein:2004:FGN

[Rot04c] Linda Rothstein. Fun and games with the NPT. Bulletin of the Atomic Scientists, 60(3):2, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/3/2.

Rothstein:2004:GWC

[Rot04d] Linda Rothstein. George W. Custer. Bulletin of the Atomic Scientists, 60(6):4, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/6/4.

Rothstein:2004:FG

[Rot04e] Linda Rothstein. So far so good? Bulletin of the Atomic Scientists, 60(5):2, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/5/2.

Rothstein:2004:SMD

[Rot04f] Linda Rothstein. Something must be done. Bulletin of the Atomic Scientists, 60(2):2, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/2/2.

Rothstein:2004:TT

[Rot04g] Linda Rothstein. A time of testing. Bulletin of the Atomic Scientists, 60(4):2, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/60/4/2.

Rothstein:2004:WSW

[Rot04h] Linda Rothstein. War in space: the wish list. Bulletin of the Atomic Scientists, 60(3):10–11, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rothstein:2005:CSA

[Rot05a] Linda Rothstein. A cover story for the ages. Bulletin of the Atomic Scientists, 61(2):4, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/2/4.

Rothstein:2005:PDP

[Rot05b] Linda Rothstein. Phantom defenses, phantom foes. Bulletin of the Atomic Scientists, 61(1):4, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/4.

Roth:2018:RCO

[Rot18] Marcie Roth. A resilient community is one that includes and protects everyone. Bulletin of the Atomic Scientists, 74(2):91– 94, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1436808.

Roush:1975:LAN

[Rou75] J. Fred Roush. Letter: Ask Norman Mailer. Bulletin of the Atomic Scientists, 31(9):4, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rowe:1977:GDJ

[Row77] Wallace P. Rowe. Guidelines that do the job: the federal rules are restrictive, but they give appropriate protection. Bulletin of the Atomic Scientists, 33(5):14–15, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sal77].

Rowen:1988:BRB

[Row88] Henry S. Rowen. Book review: American Espionage and the Soviet Target, by Jeffrey Richelson. Bulletin of the Atomic Scientists, 44(7):57–58, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Royce:1967:FFN

[Roy67] William F. Royce. Fish and fishing [Nawapa]. Bulletin of the Atomic Scientists, 23(7):26–27, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

RCEP:1976:WBG

[Roy76] Royal Commission on Environmental Pollution. A warning in Britain: Go slow on nuclear power. Bulletin of the Atomic Scientists, 32(10):22–23, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rozen:1964:SRC

[Roz64] Marvin E. Rozen. Some reflections on civil defense. Bulletin of the Atomic Scientists, 20(6):21–24, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rich:1966:HKP

[RP66] Alexander Rich and John R. Platt. How to keep the peace in a disarmed world. Bulletin of the Atomic Scientists, 22(4):14–19, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Blo66].

Rydell:1982:BWF

[RP82] Randy J. Rydell and Athanassios Platias. The Balkans a weaponfree zone? Bulletin of the Atomic Scientists, 38(5):57–59, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rudolph:1968:SRA

[RR68] Lloyd I. Rudolph and Susanne Hoeber Rudolph. Special reports from Asia: New era for India: The fourth global election. Bulletin of the Atomic Scientists, 24(2):35–40, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rajeandran:1981:EHM

[RR81a] Rajeandran and Michael R. Reich. Environmental health in Malaysia. Bulletin of the Atomic Scientists, 37(4):30–35, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ross:1981:CND

[RR81b] L. M. Ross and John R. Redick. Commentary: National disarmament mechanisms. Bulletin of the Atomic Scientists, 37(7): 60–62, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sha81a].

Ritch:1988:ACN

[RR88] John B. Ritch III and James P. Rubin. Arms control — now or never. Bulletin of the Atomic Scientists, 44(10):9–13, December 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cat89].

Ratner:2007:MOS

[RR07] Daniel Ratner and Mark A. Ratner. Make only small plans. Bulletin of the Atomic Scientists, 63(1):53–55, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/63/1/53.

Raj:2015:PRB

[RR15] Baldev Raj and P. R. Vasudeva Rao. Plutonium reprocessing, breeder reactors, and decades of debate : an Indian response. Bulletin of the Atomic Scientists, 71(4):14–17, July 2015. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1952:IGD

[RS52] Eugene Rabinowitch and John Simpson. An interview with Gordon Dean, Chairman, U.S. Atomic Energy Commission. Bulletin of the Atomic Scientists, 8(8):267–270, 274, November 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rabinowitch:1958:SE

[RS58] Eugene Rabinowitch and Mary Simpson. Science and education. Bulletin of the Atomic Scientists, 14(9):345, November 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Second author credit lost, but corrected on page 48 of November 1958 issue.

Rogers:1983:CS

[RS83] Patrick F. Rogers and Gaddis Smith. Commentary: Survivalism. Bulletin of the Atomic Scientists, 39(10):56–57, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi83].

Reicher:1988:HTP

[RS88a] Dan W. Reicher and Jason Salzman. High-tech protest against plutonium plant. Bulletin of the Atomic Scientists, 44(9):27– 30, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reicher:1988:NRP

[RS88b] Dan W. Reicher and Jason Salzman. New reactors proposed. Bulletin of the Atomic Scientists, 44(9):28, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reicher:1988:MWL

[RS88c] Dan W. Reicher and S. Jacob Scherr. Making warheads: Laying waste to the environment. Bulletin of the Atomic Scientists, 44(1):31–34, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ruehsen:2015:FPM

[RS15] Moyara Ruehsen and Leonard Spector. Follow the proliferation money. Bulletin of the Atomic Scientists, 71(5):51–58, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/71/
5/51.full.pdf+html.

Rose:2016:HSN

[RS16] Thomas Rose and Trevor Sweeting. How safe is nuclear power? A statistical study suggests less than expected. Bulletin of the Atomic Scientists, 72(2):112–115, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosenfeld:1955:FSM

[RSW55] A. H. Rosenfeld, E. J. Story, and S. D. Warshaw. Fall-out: Some measurements and damage estimates. Bulletin of the Atomic Scientists, 11(6):213–216, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Robock:2012:SAD

[RT12] Alan Robock and Owen Brian Toon. Self-assured destruction: The climate impacts of nuclear war. Bulletin of the Atomic Scientists, 68(5):66–74, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/68/5/66.

Robock:2019:HIP

[RTB^{+}19] Alan Robock, Owen B. Toon, Charles G. Bardeen, Lili Xia, Hans M. Kristensen, Matthew McKinzie, R. J. Peterson, Cheryl S. Harrison, Nicole S. Lovenduski, and Richard P. Turco. How an India–Pakistan nuclear war could start — and have global consequences. Bulletin of the Atomic Scientists, 75(6): 273–279, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1680049.

Ruby:1962:ECA

[Rub62] Lawrence Ruby. Events and comments: Atmospheric testing. Bulletin of the Atomic Scientists, 18(1):31, January 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubel:1963:MS

[Rub63] John H. Rubel. The military in space. Bulletin of the Atomic Scientists, 19(5):20–22, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubinstein:1965:FIF

[Rub65] Alvin Z. Rubinstein. Forum: On IAEA’s future. Bulletin of the Atomic Scientists, 21(1):25–27, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubin:1970:AG

[Rub70] Morton J. Rubin. Antarctic geology. Bulletin of the Atomic Scientists, 26(10):48–54, December 1970. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Ruby:1973:BBE

[Rub73] Lawrence Ruby. Books: Nuclear Energy — Its Physics and Its Social Challenge, by D. R. Inglis. Bulletin of the Atomic Scientists, 29(7):52–53, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubin:1974:PES

[Rub74] Milton D. Rubin. Plugging the energy sieve: More efficient utilization would go far toward meeting the crisis: Here’s what we have to do. Bulletin of the Atomic Scientists, 30(10):7–17, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubin:1987:SDR

[Rub87] James P. Rubin. The superpower dispute over radars. Bulletin of the Atomic Scientists, 43(3):34–37, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubin:1990:IBM

[Rub90] Uzi Rubin. Iraq and the ballistic missile scare. Bulletin of the Atomic Scientists, 46(8):11–13, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rubinson:2008:CSU

[Rub08] Paul Harold Rubinson. Containing Science: The U.S. National Security State and Scientists’ Challenge to Nuclear Weapons during the Cold War. Ph.D. thesis, The University of Texas at Austin, Austin, TX, USA, August 2008. xi + 445 pp. URL

https://www.lib.utexas.edu/etd/d/2008/rubinsonp66913/
rubinsonp66913.pdf.

Ruff:1990:BTA

[Ruf90] Tilman Ruff. Bomb tests attack the food chain. Bulletin of the Atomic Scientists, 46(2):32–34, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ruina:1968:NAR

[Rui68] J. P. Ruina. The nuclear arms race: Diagnosis and treatment. Bulletin of the Atomic Scientists, 24(8):19–22, October 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rumsey:1954:LEM

[Rum54] C. H. Rumsey family. Letter to the Editor: Madame Joliot and the American Chemical Society. Bulletin of the Atomic Scientists, 10(6):212, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rummerfield:1971:CCC

[Rum71a] Philip S. Rummerfield. Communications: Climatic changes. Bulletin of the Atomic Scientists, 27(3):2–3, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pec70].

Rummerfield:1971:CRE

[Rum71b] Philip S. Rummerfield. Communications: Radiation effects controversy. Bulletin of the Atomic Scientists, 27(2):3–4, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TG70].

Russell:1946:ABP

[Rus46] Bertrand Russell. The atomic bomb and the prevention of war. Bulletin of the Atomic Scientists, 2(7–8):19–21, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rush:1950:LEF

[Rus50] J. H. Rush. Letter to the Editor: Fusion bombs. Bulletin of the Atomic Scientists, 6(5):138, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rush:1951:BBL

[Rus51a] J. H. Rush. Books: Science and Loyalty, by Walter Gellhorn, and The Tenney Committee, by Edward L. Barrett, Jr. Bulletin of the Atomic Scientists, 7(11):346–347, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1951:RM

[Rus51b] Bertrand Russell. The root of the matter. Bulletin of the Atomic Scientists, 7(7–8):203, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1952:LEL

[Rus52a] Bertrand Russell. “The love of excitement … ”. Bulletin of the Atomic Scientists, 8(6):199, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1952:NPO

[Rus52b] Henry Norris Russell. “New powers are in our hands”. Bulletin of the Atomic Scientists, 8(2):35, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1953:WWH

[Rus53] Bertrand Russell. “What would help mankind most”. Bulletin of the Atomic Scientists, 9(10):379, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpted from the author’s article of the same title, published in the New York Times, 27 September 1953.

Russell:1954:DM

[Rus54] Bertrand Russell. The danger to mankind. Bulletin of the Atomic Scientists, 10(1):8–9, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1955:CC

[Rus55a] Bertrand Russell. Current comment. Bulletin of the Atomic Scientists, 11(2):67–68, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the author’s article Commission for Peace in the 18 December 1954 issue of The Nation.

Russell:1955:OS

[Rus55b] Bertrand Russell. The optimism of science. Bulletin of the Atomic Scientists, 11(3):75, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1955:SNY

[Rus55c] Bertrand Russell. A statement for the New Year. Bulletin of the Atomic Scientists, 11(1):4, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1956:WCS

[Rus56a] Lord Bertrand Russell. World conference of scientists. Bulletin of the Atomic Scientists, 12(2):41–43, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1956:RMG

[Rus56b] W. L. Russell. Radiation in mice — the genetic effects and their implications for man. Bulletin of the Atomic Scientists, 12(1): 19–20, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1958:CPA

[Rus58a] Bertrand Russell. The case for Pugwash: An answer to Mr. Marseille. Bulletin of the Atomic Scientists, 14(4):144–146, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Mar58c].

Russell:1958:OWG

[Rus58b] Bertrand Russell. Only world government can prevent the war nobody can win. Bulletin of the Atomic Scientists, 14(7):259– 261, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1960:SRS

[Rus60] Bertrand Russell. The social responsibilities of the scientist. Bulletin of the Atomic Scientists, 16(6):200, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Twoparagraph excerpt for the author’s article in Science, 12 February 1960.

Russell:1962:CBN

[Rus62a] Bertrand Russell. The case for British nuclear disarmament. Bulletin of the Atomic Scientists, 18(3):6–10, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rust:1962:BBD

[Rus62b] John H. Rust. Books: The Dreams of Reasons. Science and Utopias, by Rene DuBos. Bulletin of the Atomic Scientists, 18(1): 40–41, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russett:1970:CLC

[Rus70a] Bruce M. Russett. Comment: Licensing: For cars and babies. Bulletin of the Atomic Scientists, 26(9):15–19, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HG71] and reply [Rus71].

Russett:1970:VRA

[Rus70b] Bruce M. Russett. Vietnam and restraints on aerial warfare. Bulletin of the Atomic Scientists, 26(1):9–12, January 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russett:1971:CRR

[Rus71] Bruce M. Russett. Communications: Russett replies. Bulletin of the Atomic Scientists, 27(5):3, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rus70a, HG71].

Russell:1978:DPC

[Rus78] Bertrand Russell. Dear Pugwash conferees. Bulletin of the Atomic Scientists, 34(4):36–37, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Russell:1992:LNT

[Rus92a] Ruth B. Russell. Letter: No threats. Bulletin of the Atomic Scientists, 48(1):45, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HPR91].

Russell:1992:LRU

[Rus92b] Ruth B. Russell. Letter: Reforming U.N. finances. Bulletin of the Atomic Scientists, 48(8):46–47, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch92b].

Russell:2020:OWG

[Rus20] Bertrand Russell. 1958: Only world government can prevent the war nobody can win. Bulletin of the Atomic Scientists, 76 (6):359–362, 2020. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2020.1847482.

Rich:1964:FAD

[RV64] Alexander Rich and Aleksandr P. Vinogradov. Forum: Arctic disarmament. Bulletin of the Atomic Scientists, 20(9):22–23, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ross:1976:EEO

[RW76] Marc H. Ross and Robert H. Williams. Energy efficiency: our most underrated energy resource. Bulletin of the Atomic Scientists, 32(9):30–38, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kre77].

Ross:1980:PFC

[RW80] Marc H. Ross and Robert H. Williams. Perspective: A fuel conservation strategy. Bulletin of the Atomic Scientists, 36(9):1, 56, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hai81].

Wei:1947:PPS

[rWeKHF47] Hsioh ren Wei, Jamal el Kourdajy, Ralph L. Harry, and Paul C. Fine. Processing and purification of source materials. Bulletin of the Atomic Scientists, 3(9):259, 265, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Reagan:1983:OUS

[RWK^{+}83] President Ronald Reagan, Caspar Weinberger, George Keyworth, John W. Vessey, Jr., George Milburn, Daniel O. Graham, Kosta Tsipis, William J. Perry, Harold Brown, Wolfgang K. H. Panofsky, Kurt Gottfried, George Rathjens, and Jack Ruina. Onward and upward with space defense. Bulletin of the Atomic Scientists, 39(6):4–8, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ryan:1989:LCS

[Rya89] Daniel Ryan. Letter: Clock stopped World War III. Bulletin of the Atomic Scientists, 45(10):49, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ryan:2007:SF

[Rya07] Missy Ryan. States of failure. Bulletin of the Atomic Scientists, 63(3):??, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/3/50.

Rybeck:1971:COO

[Ryb71] S. Arthur Rybeck, Jr. Communications: One’s own species. Bulletin of the Atomic Scientists, 27(3):2, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lor70].

S:1946:SCA

[S.46] E. S. A Soviet comment on American atomic policy. Bulletin of the Atomic Scientists, 1(12):16, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

S:1955:APU

[S.55a] M. S. Atoms for peace in the U.N. Bulletin of the Atomic Scientists, 11(1):24–27, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

S:1955:CCS

[S.55b] M. S. Current comment: The strategy of nuclear war. Bulletin of the Atomic Scientists, 11(3):100–101, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

S:1980:RTB

[S.80] B. A. S. The reading table: Iran: From Religious Dispute to Revolution, by Michael M. J. Fischer; The New Authoritarianism in Latin America, edited by David Collier. Bulletin of the Atomic Scientists, 36(7):46–47, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1998:EBA

[S^{+}98] Stephen I. Schwartz et al. Excerpts from Atomic Audit. Bulletin of the Atomic Scientists, 54(5):36–43, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pen98].

Seaborg:1962:NPN

[SA62] Glenn T. Seaborg and Samuel K. Allison. Nuclear power and the next twenty years. Bulletin of the Atomic Scientists, 18(10):12– 14, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Speed:1982:CMS

[SA82] Roger D. Speed and J. Edward Anderson. Commentary: Minuteman survivability. Bulletin of the Atomic Scientists, 38(4):52–53, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [And81].

Saaty:1984:BND

[Saa84] Thomas L. Saaty. Breaking the nuclear deadlock: Center for conflict resolution. Bulletin of the Atomic Scientists, 40(3):52, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sabin:1963:ICB

[Sab63] Albert B. Sabin. International cooperation: Breaking the impasse. Bulletin of the Atomic Scientists, 19(8):36, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sabin:1972:CAS

[Sab72] Albert B. Sabin. Conditions for the application of science and technology for human welfare. Bulletin of the Atomic Scientists, 28(8):7–30, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Sabo:1976:LGC

[Sab76] Susan J. Sabo. Letter: Guilty consciences. Bulletin of the Atomic Scientists, 32(3):2, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shultz:2016:FFU

[SAB^{+}16] James M. Shultz, Benjamin M. Althouse, Florence Baingana, Janice L. Cooper, Maria Espinola, M. Claire Greene, Zelde Espinel, Clyde B. McCoy, Laurie Mazurik, and Andreas Rechkemmer. Fear factor: The unseen perils of the Ebola outbreak. Bulletin of the Atomic Scientists, 72(5):304–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sacher:1946:IBS

[Sac46] G. Sacher. International Bureau of Standards. Bulletin of the Atomic Scientists, 1(10):12, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sacher:1948:MR

[Sac48] George Sacher. Manpower for research. Bulletin of the Atomic Scientists, 4(2):57–58, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sachs:1957:LEC

[Sac57] Albert P. Sachs. Letter to the Editor: The contribution of science to religion. Bulletin of the Atomic Scientists, 13(4):153, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor57b].

Sachs:1964:PPE

[Sac64] Robert G. Sachs. Power of prediction — an example. Bulletin of the Atomic Scientists, 20(10):20–21, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sachs:1972:CRM

[Sac72a] Leo Sachs. Cancer research: A molecular approach to therapy. Bulletin of the Atomic Scientists, 28(8):52–55, October 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Sachs:1972:BBS

[Sac72b] Robert G. Sachs. Books: A Special Interest, by Leonard Greenbaum. Bulletin of the Atomic Scientists, 28(6):51–55, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gre72].

Sachs:1984:BRB

[Sac84] Robert G. Sachs. Book review: Nuclear Power in an Age of Uncertainty. Bulletin of the Atomic Scientists, 40(8):32–36, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sadron:1952:SFE

[Sad52] Charles Sadron. Some French experiences. Bulletin of the Atomic Scientists, 8(7):244–245, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Sagittarius:1950:GA

[Sag50] Sagittarius. Galluping atom. Bulletin of the Atomic Scientists, 6(10):292, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Six-stanza poem.

Sagan:1967:UFO

[Sag67] Carl Sagan. Unidentified flying objects. Bulletin of the Atomic Scientists, 23(6):43–44, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar67a] and comment [Mun67a].

Sagan:1969:CIM

[Sag69] Leonard A. Sagan. Comments: Infant mortality controversy: Sternglass and his critics. Bulletin of the Atomic Scientists, 25(8): 26–28, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e, Wei69b, Wor69, Ste69c].

Sagan:1971:CH

[Sag71] Ginetta Sagan. Communications: On Hiroshima. Bulletin of the Atomic Scientists, 27(2):2–3, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shr70].

Sagan:1973:SEH

[Sag73] Carl Sagan. Space exploration as a human enterprise: The scientific interest. Bulletin of the Atomic Scientists, 29(9):30–33, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi74c].

Sagan:1974:MM

[Sag74] Carl Sagan. The mountains of Mars. Bulletin of the Atomic Scientists, 30(3):42–46, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sagan:1983:PWG

[Sag83] Carl Sagan. To preserve a world graced by life. Bulletin of the Atomic Scientists, 39(1):2–3, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sagasti:1986:BRB

[Sag86] Francisco R. Sagasti. Book review: Technology Transfer and Human Values: Concepts, Applications, Cases, edited by Peter B. Heller. Bulletin of the Atomic Scientists, 42(5):43–44, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sagan:1992:BE

[Sag92] Carl Sagan. Between enemies. Bulletin of the Atomic Scientists, 48(4):24–26, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sagdeev:1993:RSS

[Sag93] Roald Sagdeev. Russian scientists save American secrets. Bulletin of the Atomic Scientists, 49(4):32–36, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saha:2020:CRC

[Sah20] Sagatom Saha. The climate risks of China’s Belt and Road Initiative. Bulletin of the Atomic Scientists, 76(5):249–255, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1806584.

Sailor:1972:RLH

[Sai72] Vance L. Sailor. The role of the Lloyd Harbor Study Group in the Shoreham hearings — an assessment. Bulletin of the Atomic Scientists, 28(6):25–31, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [Car72, Sch72a].

Sailor:1998:LDR

[Sai98] William C. Sailor. Letter: Don’t reprocess. Bulletin of the Atomic Scientists, 54(3):4, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98].

Saideh:2007:UP

[Sai07] Lotfian Saideh. An unnecessary pursuit. Bulletin of the Atomic Scientists, 63(4):35, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/4/35.

Sailor:2009:NNA

[Sai09] William C. Sailor. A nonproliferation network at Los Alamos? Bulletin of the Atomic Scientists, 65(3):34–39, May 2009. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/65/3/34.

Sailor:2010:CUN

[Sai10] William C. Sailor. Creating the ultimate nuclear reactor. Bulletin of the Atomic Scientists, 66(4):23–32, July/August 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/4/23.

Saif:2014:ZCW

[Sai14] Mostafa Elwi Saif. In the zone? Chemical weapons and the Middle East: The Egyptian response. Bulletin of the Atomic Scientists, 70(3):1–3, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakurai:1969:BBH

[Sak69] Jun John Sakurai. Books: A History of Japanese Astronomy — Chinese Background and Western Impact, by Shigeru Nakayama. Bulletin of the Atomic Scientists, 25(10):38, 40, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakharov:1975:NOW

[Sak75] Andrei Sakharov. Need for an open world. Bulletin of the Atomic Scientists, 31(9):8–9, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakharov:1978:NEF

[Sak78] Andrei D. Sakharov. Nuclear energy and the freedom of the West. Bulletin of the Atomic Scientists, 34(6):12–14, June 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wou78].

Sakharov:1979:CFP

[Sak79] Andrei Sakharov. Commentary: The first priority: disarmament. Bulletin of the Atomic Scientists, 35(5):9, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakharov:1982:BSL

[Sak82] Andrei Sakharov. Bulletin: The Sakharov letter. Bulletin of the Atomic Scientists, 38(9):61–62, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakharov:1983:MG

[Sak83] Andrei Sakharov. A message from Gorky. Bulletin of the Atomic Scientists, 39(6):2–3, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakamoto:1984:MPR

[Sak84] Yoshikazu Sakamoto. Major power relations in East Asia. Bulletin of the Atomic Scientists, 40(2):19–24, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakamoto:1988:ASJ

[Sak88] Yoshikazu Sakamoto. After the summit: Japanese skeptical of Eurocentric accord. Bulletin of the Atomic Scientists, 44(2):30– 32, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sakamoto:1998:LNA

[Sak98] Yoshikazu Sakamoto. Letter: Needed: activist citizens. Bulletin of the Atomic Scientists, 54(5):67–68, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salisbury:1953:SMS

[Sal53] Morse Salisbury. Statement by Morse Salisbury, Directory, Division of Information Services, USAEC. Bulletin of the Atomic Scientists, 9(3):75, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bor53].

Salam:1963:DRD

[Sal63a] Abdus Salam. Diseases of the rich and diseases of the poor. Bulletin of the Atomic Scientists, 19(4):3, 27, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salisbury:1963:BBT

[Sal63b] Harrison E. Salisbury. Books: The Two Faces of Tass, by Theodore E. Kruglak. Bulletin of the Atomic Scientists, 19(2): 33–34, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1964:PCT

[Sal64a] Abdus Salam. Pakistan: The case for technological development. Bulletin of the Atomic Scientists, 20(3):2–5, March 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1964:WSD

[Sal64b] Abdus Salam. World security and developing nations. Bulletin of the Atomic Scientists, 20(7):6–8, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1965:NCP

[Sal65] Abdus Salam. A new center for physics. Bulletin of the Atomic Scientists, 21(10):43–45, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1968:UNI

[Sal68] Abdus Salam. The United Nations and the international world of physics. Bulletin of the Atomic Scientists, 24(2):14–15, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salloch:1969:CMM

[Sal69] Roger Salloch. Cambridge: March 4, the movement, and M.I.T. Bulletin of the Atomic Scientists, 25(5):32–35, May 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1970:MWU

[Sal70] Abdus Salam. Memorandum on a world university. Bulletin of the Atomic Scientists, 26(3):38–39, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salpeter:1971:BBM

[Sal71] Edwin E. Salpeter. Books: The Making of an Ex-Astronaut, by Brian O’Leary. Bulletin of the Atomic Scientists, 27(3):44, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1973:UNU

[Sal73] Abdus Salam. The United Nations University. Bulletin of the Atomic Scientists, 29(7):41, 56, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salki:1975:LNA

[Sal75a] Djozic Salki. Letters: Non-aligned. Bulletin of the Atomic Scientists, 31(8):3, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch75c].

Salzman:1975:LSE

[Sal75b] Lorna Salzman. Letter: Scientific elite. Bulletin of the Atomic Scientists, 31(9):2–3, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Pri76b, Fes76, Maw76].

Salam:1976:IR

[Sal76] Abdus Salam. Ideals and realities. Bulletin of the Atomic Scientists, 32(7):9–15, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gla76c, Sto76b, Rao77].

Salzman:1977:LDU

[Sal77] Lorna Salzman. Letter: DNA: The ultimate question. Bulletin of the Atomic Scientists, 33(8):4, 6–7, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Row77].

Salaff:1978:LD

[Sal78a] Stephen Salaff. The Lucky Dragon. Bulletin of the Atomic Scientists, 34(5):21–23, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Illustrations by Ben Shahn.

Salisbury:1978:LMF

[Sal78b] Henry M. Salisbury. Letter: Mutual fears. Bulletin of the Atomic Scientists, 34(2):7, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salisbury:1979:PHN

[Sal79a] Harrison Salisbury. Panel I: How a nuclear war can start … in the Far East. Bulletin of the Atomic Scientists, 35(4):22, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salisbury:1979:LBC

[Sal79b] Henry M. Salisbury. Letter: the Bulletin’s clock. Bulletin of the Atomic Scientists, 35(3):88, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salaff:1980:CHH

[Sal80a] Stephen Salaff. Commentary: Help for the Hibakusha. Bulletin of the Atomic Scientists, 36(4):61–63, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salaff:1980:PCE

[Sal80b] Stephen Salaff. The plutonium connection: energy and arms. Bulletin of the Atomic Scientists, 36(7):18–20, September 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1980:TTR

[Sal80c] Abdus Salam. From Toledo to Trieste — renewing our commitment. Bulletin of the Atomic Scientists, 36(7):35–38, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salzman:1987:GNZ

[Sal87] Jason Salzman. The genesis of New Zealand’s ban. Bulletin of the Atomic Scientists, 43(6):45–46, 48–49, July/August 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salam:1988:PWT

[Sal88] Abdus Salam. Perspective: What the Third World really needs. Bulletin of the Atomic Scientists, 44(9):8–10, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saleska:1990:LLR

[Sal90a] Scott Saleska. Low-level radioactive waste: Gamma rays in the garbage. Bulletin of the Atomic Scientists, 46(3):18–25, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saleska:1990:MRE

[Sal90b] Scott Saleska. Measures of radioactivity and exposure. Bulletin of the Atomic Scientists, 46(3):21, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saleska:1990:WIR

[Sal90c] Scott Saleska. What is ionizing radiation? Bulletin of the Atomic Scientists, 46(3):20, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salzman:2016:WCC

[Sal16] Rachel S. Salzman. Will climate-change efforts affect EU–Russian relations? (Probably not). Bulletin of the Atomic Scientists, 72 (6):384–389, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Samios:1986:BS

[Sam86] Nick P. Samios. Brookhaven and SDI. Bulletin of the Atomic Scientists, 42(3):56–57, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol86b] and comment [Wol86].

Samson:2006:HD

[Sam06] Victoria Samson. Hacker’s delight. Bulletin of the Atomic Scientists, 62(5):64–65, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/64.

Samore:2018:NKV

[Sam18] Gary Samore. North Korean verification: Good enough for government work? Bulletin of the Atomic Scientists, 74(5): 312–316, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1507787.

Samson:2022:CRP

[Sam22] Victoria Samson. The complicating role of the private sector in space. Bulletin of the Atomic Scientists, 78(1):6–10, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.2014229.

Sandys:1957:BWP

[San57] Duncan Sandys. British white paper on defense. Bulletin of the Atomic Scientists, 13(6):223, 232, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sands:1963:MTB

[San63] Matthew Sands. Monitoring a test ban. Bulletin of the Atomic Scientists, 19(3):12–18, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sanders:1966:APO

[San66a] Irwin T. Sanders. American professionals overseas. Bulletin of the Atomic Scientists, 22(10):40–45, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sanders:1966:APS

[San66b] Ralph Sanders. The autumn of power: The scientist in the political establishment. Bulletin of the Atomic Scientists, 22(8): 22–25, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sandler:1968:WC

[San68] Åke Sandler. A world constitution? Bulletin of the Atomic Scientists, 24(4):38, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sandidge:1976:LDB

[San76a] Ernie L. Sandidge. Letter: Dropping the bomb. Bulletin of the Atomic Scientists, 32(2):3, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber75b].

Sandoval:1976:CPA

[San76b] R. Robert Sandoval. Consider the porcupine: another view of nuclear proliferation. Bulletin of the Atomic Scientists, 32(5): 17–23, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ber76a].

Sandor:1979:BC

[San79] David Sandor. That Progressive case. Bulletin of the Atomic Scientists, 35(7):64, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79i].

Sanders:1990:NRN

[San90] Ben Sanders. NPT review: Non-proliferation treaty: a broken record? Bulletin of the Atomic Scientists, 46(6):15–18, July/ August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sanders:1991:RNN

[San91] Ben Sanders. Reports: NPT: North Korea, South Africa ready to tell all? Bulletin of the Atomic Scientists, 47(7):8–10, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sankaran:2014:DPD

[San14] Jaganath Sankaran. Destroying Pakistan to deter India? The problem with Pakistan’s battlefield nukes. Bulletin of the Atomic Scientists, 70(4):74–84, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Santos:2016:BRA

[San16] Paulo E. Santos. To ban or regulate autonomous weapons: A Brazilian response. Bulletin of the Atomic Scientists, 72(2):117– 119, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sapolsky:1968:ANF

[Sap68] Harvey M. Sapolsky. Advisory network for Federal service. Bulletin of the Atomic Scientists, 24(3):46–48, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saperstein:1975:ENN

[Sap75a] Alvin M. Saperstein. Editorial: On the need for a new acquisition. Bulletin of the Atomic Scientists, 31(6):10, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saperstein:1975:LCB

[Sap75b] Alvin M. Saperstein. Letter: Cart before the horse. Bulletin of the Atomic Scientists, 31(9):3, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saperstein:1980:CNE

[Sap80a] Alvin M. Saperstein. Commentary: A non-economist looks at inflation. Bulletin of the Atomic Scientists, 36(5):61–63, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saperstein:1980:CFP

[Sap80b] Alvin M. Saperstein. Commentary: ‘First to play cannot lose’. Bulletin of the Atomic Scientists, 36(1):63–64, January 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saperstein:1984:CNF

[Sap84] Alvin S. Saperstein. Commentary: No first use: yet another view from Europe. Bulletin of the Atomic Scientists, 40(1):62– 63, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TC83].

Saperstein:1988:SMC

[Sap88] Alvin M. Saperstein. SDI a model for chaos. Bulletin of the Atomic Scientists, 44(8):40–43, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sarabhai:1972:IGR

[Sar72] Vikram Sarabhai. India and the Green Revolution. Bulletin of the Atomic Scientists, 28(7):8–10, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sartori:1978:BBD

[Sar78] L. Sartori. Books: Strategic Disarmament, Verification and National Security, by Andrzej Karkoszka. Bulletin of the Atomic Scientists, 34(6):55–58, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sartori:1983:WTP

[Sar83] Leo Sartori. The weapons tutorial — Part 5: When the bomb falls. Bulletin of the Atomic Scientists, 39(6):40–47, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Parts 1 [Mor83c], 2 [Tsi83i], 3 [Mar83b], 4 [Hol83c], and 6 [Rom83].

Sartori:1985:BRB

[Sar85] Leo Sartori. Book review: From H-Bombs to Star Wars: The Politics of Strategic Decision Making, by Jonathan B. Stein. Bulletin of the Atomic Scientists, 41(3):57–58, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saroyan:1989:TT

[Sar89] Mark Saroyan. Trouble in the Transcaucasus. Bulletin of the Atomic Scientists, 45(2):16–18, 20, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sarkar:2017:MNR

[Sar17] Jayita Sarkar. Managing nuclear risk in South Asia. Bulletin of the Atomic Scientists, 73(1):59–61, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2016.
1264215.

Sass:2006:NSP

[Sas06] Jennifer Sass. No small problem. Bulletin of the Atomic Scientists, 62(2):23–24, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/2/23.

Sasikumar:2017:ANM

[Sas17] Karthika Sasikumar. After nuclear midnight: The impact of a nuclear war on India and Pakistan. Bulletin of the Atomic Scientists, 73(4):226–232, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1338009.

Sauberman:1974:LHS

[Sau74] N. H. Sauberman. Letters: He’s scared, too. Bulletin of the Atomic Scientists, 30(6):6–7, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gou74].

Sauberman:1975:LCL

[Sau75] Nat. H. Sauberman. Letters: Credibility lost. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Sauberman:1978:LFD

[Sau78] N. H. Sauberman. Letter: Facing the dilemma. Bulletin of the Atomic Scientists, 34(4):4–5, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Par78b, Par78a].

Sauer:2001:MDW

[Sau01] Tom Sauer. Missile defense: Wrong in too many ways. Bulletin of the Atomic Scientists, 57(6):36–38, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/6/36.

Sauer:2010:UTN

[Sau10] Tom Sauer. U.S. tactical nuclear weapons: a European perspective. Bulletin of the Atomic Scientists, 66(5):65–75, September/ October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
5/65.

Sauer:2017:HWN

[Sau17] Tom Sauer. How will NATO’s non-nuclear members handle the UN’s ban on nuclear weapons? Bulletin of the Atomic Scientists, 73(3):177–181, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1315039.

Sawyer:1953:YMP

[Saw53] Roland Sawyer. It’s up to you, Mr. President. Bulletin of the Atomic Scientists, 9(7):244–246, 272, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sawyer:1954:HBC

[Saw54a] Roland Sawyer. The H-bomb chronology. Bulletin of the Atomic Scientists, 10(7):287–290, 300, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sawyer:1954:MTS

[Saw54b] Roland Sawyer. More than security. Bulletin of the Atomic Scientists, 10(7):284, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the author’s article in The Christian Science Monitor.

Sawyer:1963:HCP

[Saw63] Jack Sawyer. How can psychology contribute? Bulletin of the Atomic Scientists, 19(9):35–36, 38, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sawyier:1979:BRB

[Saw79] Fay Horton Sawyier. Book review: Morals, Science and Sociality, edited by H. Tristram Engelhardt, Jr., and Daniel Callahan. Bulletin of the Atomic Scientists, 35(10):41–42, December 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sawyier:1980:BRB

[Saw80] Fay Horton Sawyier. Book review: The Limits of Scientific Inquiry, by G. Holton and R. S. Morison. Bulletin of the Atomic Scientists, 36(4):48, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sawyier:1981:BRBb

[Saw81a] Fay Horton Sawyier. Book review: Ethics in an Age of Pervasive Technology, by Melvin Kranzberg. Bulletin of the Atomic Scientists, 37(8):41, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sawyier:1981:BRBa

[Saw81b] Fay Horton Sawyier. Book review: The Tanner Lecture on Human Values, edited by Sterling McMurin. Bulletin of the Atomic Scientists, 37(4):49, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sax:1949:CAG

[Sax49] Karl Sax. Commentary by American geneticists: Genetics and agriculture. Bulletin of the Atomic Scientists, 5(5):143, 146, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sax:1951:FRP

[Sax51] Karl Sax. Food resources and population growth. Bulletin of the Atomic Scientists, 7(4):105–107, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sayre:1955:SMH

[Say55] Anne Sayre. The scientific method in human affairs. Bulletin of the Atomic Scientists, 11(8):295–296, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kir55].

Sayer:1984:FRB

[Say84] Karen Sayer. Film reviews: In the Nuclear Shadow: What Can the Children Tell Us?, produced by Eric Thiermann, Vivienne Verdon-Roe and Ian Thiermann; There’s a Nuclear War Going on Inside Me; Third Graders Discuss Nuclear Issues; Fifth Graders Discuss Nuclear Issues; Ninth Graders Discuss Nuclear Issues; Bombs Will Make the Rainbow Break. Bulletin of the Atomic Scientists, 40(5):48–49, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sayles:1992:LBL

[Say92] Daniel Sayles. Letter: Bulletin a limp noodle. Bulletin of the Atomic Scientists, 48(10):51, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sayigh:1993:LID

[Say93] Yezid Sayigh. Letter: Inherently destabilizing. Bulletin of the Atomic Scientists, 49(8):53–54, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coh93a].

Steele:1973:CPP

[SB73] James H. Steele and George M. Baer. Communications: Planned parenthood for pets. Bulletin of the Atomic Scientists, 29(7): 54, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DIJ73, Dje73].

Schmidt:1977:LPM

[SB77] F. H. Schmidt and D. Bodansky. Letter: Pseudo-McCarthyism. Bulletin of the Atomic Scientists, 33(3):2, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76a].

Smit:1980:LF

[SB80] Wim A. Smit and Peter Boskma. Laser fusion. Bulletin of the Atomic Scientists, 36(10):34–38, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1984:THB

[SB84] Nancy Smith and Barton J. Bernstein. Truman and the H-bomb. Bulletin of the Atomic Scientists, 40(8):44–47, October 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber84c].

Scheiner:1989:LAC

[SB89] Charles A. Scheiner and Nathaniel S. Borenstein. Letters: All collaboration wrong. Bulletin of the Atomic Scientists, 45(6):54, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bor89].

Stanford:1991:BRP

[SB91] George S. Stanford and Robert E. Boyar. Book review: Power players: Nuclear Politics, by James M. Jasper, Princeton University Press, 1991. Bulletin of the Atomic Scientists, 47(4):45–46, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jas91].

Sullivan:2023:NFN

[SB23] Kathleen Sullivan and Matthew Breay Bolton. Nuclear-free NYC: How New Yorkers are disarming the legacies of the Manhattan Project. Bulletin of the Atomic Scientists, 80(1):32–37, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2293573.

Sperlich:1975:LPV

[SC75] Harold K. Sperlich and James Cicarelli. Letters: Pistons vs. turbines. Bulletin of the Atomic Scientists, 31(5):3–4, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cic74].

Simens:1976:LRH

[SC76] H. G. Simens and David Dinsmore Comey. Letter: Red herring. Bulletin of the Atomic Scientists, 32(2):64, February 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com75e].

Shore:1997:LTO

[SC97] Lloyd G. Shore and Alan Cranston. Letter: “Targets of Opportunity” hits bull’s-eye. Bulletin of the Atomic Scientists, 53(6): 3, 62, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kri97].

Sokova:2007:NPB

[SC07] Elena Sokova and Cristina Hansell Chuen. Nuclear power broker. Bulletin of the Atomic Scientists, 63(5):50–54, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/63/5/50.

Scarlott:1988:UOP

[Sca88] Jennifer Scarlott. U.S. offers Palau dollars or democracy. Bulletin of the Atomic Scientists, 44(9):31–35, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scarlott:1991:BRP

[Sca91] Jennifer Scarlott. Book review: “Protecting” Palau: Overreaching in Paradise: United States Policy in Palau Since 1945, by Sue Rabbitt Roff. Bulletin of the Atomic Scientists, 47(3):44–45, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scambos:2011:EIS

[Sca11] Ted Scambos. Earth’s ice: Sea level, climate, and our future commitment. Bulletin of the Atomic Scientists, 67(1):28–40, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/1/28.

Scarry:2014:FW

[Sca14] Elaine Scarry. The floor of the world. Bulletin of the Atomic Scientists, 70(2):21–35, March 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schurr:1947:EAA

[Sch47] Sam H. Schurr. Economic aspects of atomic energy as a source of power: A discussions of the estimated costs of atomic power and of the economic factors which may affect its use. Bulletin of the Atomic Scientists, 3(4–5):117–120, 125, April/May 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schurr:1949:APS

[Sch49] Sam H. Schurr. Atomic power in selected industries. Bulletin of the Atomic Scientists, 5(11):301–308, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1952:SFE

[Sch52] Lawrence Schwartz. Some French experiences. Bulletin of the Atomic Scientists, 8(7):237–238, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Schein:1954:BBM

[Sch54a] Edgar H. Schein. Books: The Making of a Scientist, by Anne Roe. Bulletin of the Atomic Scientists, 10(7):297–298, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano54k].

Schlem:1954:LEP

[Sch54b] R. Schlem. Letter to the Editor: Population planning. Bulletin of the Atomic Scientists, 10(8):337, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1954:SAL

[Sch54c] Harry Schwartz. Soviet atom lead overcome by U.S. Bulletin of the Atomic Scientists, 10(5):170, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schweitzer:1954:SMS

[Sch54d] Albert Schweitzer. The scientists must speak up. Bulletin of the Atomic Scientists, 10(9):339, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schmitt:1957:LET

[Sch57a] Roland W. Schmitt. Letter to the Editor: The telescope and the church. Bulletin of the Atomic Scientists, 13(5):184, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schweitzer:1957:AEN

[Sch57b] Albert Schweitzer. Appeal to end nuclear tests. Bulletin of the Atomic Scientists, 13(6):204–205, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Lib57].

Schenck:1958:ISE

[Sch58a] Hubert G. Schenck. Impact of science in East Asia. Bulletin of the Atomic Scientists, 14(7):273–275, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwab:1958:TSI

[Sch58b] Joseph J. Schwab. The teaching of science as inquiry. Bulletin of the Atomic Scientists, 14(9):374–379, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schelling:1959:SAD

[Sch59a] Thomas C. Schelling. Surprise attack and disarmament. Bulletin of the Atomic Scientists, 15(10):413–418, December 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schilpp:1959:DPA

[Sch59b] Paul Arthur Schilpp. Does philosophy have anything to say to our age? Bulletin of the Atomic Scientists, 15(5):216–218, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schubert:1959:FIF

[Sch59c] Jack Schubert. Fetal irradiation and fallout. Bulletin of the Atomic Scientists, 15(6):253–256, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schelling:1960:MMW

[Sch60a] Thomas C. Schelling. Meteors, mischief, and war. Bulletin of the Atomic Scientists, 16(7):292–296, 300, September 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schelling:1960:RSI

[Sch60b] Thomas C. Schelling. The retarded science of international strategy. Bulletin of the Atomic Scientists, 16(3):103–106, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schick:1962:ILO

[Sch62a] F. B. Schick. International law in outer space. Bulletin of the Atomic Scientists, 18(9):2–6, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schiller:1962:ARM

[Sch62b] Herbert Schiller. Access to raw materials. Bulletin of the Atomic Scientists, 18(8):16–19, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1962:URS

[Sch62c] Leonard E. Schwartz. U.N. resumes space cooperation efforts. Bulletin of the Atomic Scientists, 18(7):44–45, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schick:1963:NRA

[Sch63a] F. B. Schick. News & reports: Aerospace in the Nuclear Age. Bulletin of the Atomic Scientists, 19(10):46–49, December 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schrank:1963:LEO

[Sch63b] Alice DeGanton Schrank. Letter to the Editor: Open letter to scientists. Bulletin of the Atomic Scientists, 19(4):34–35, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schuster:1963:BBS

[Sch63c] Richard P. Schuster, Jr. Books: Conflict in Space, by M. N. Golovine. Bulletin of the Atomic Scientists, 19(5):32–33, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schuster:1963:LEI

[Sch63d] Richard P. Schuster, Jr. Letter to the Editor: Information for arms control. Bulletin of the Atomic Scientists, 19(2):35–36, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1963:WIS

[Sch63e] Leonard E. Schwartz. When is international space cooperation international? Bulletin of the Atomic Scientists, 19(6):12–18, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schelling:1965:SFA

[Sch65a] T. C. Schelling. Signals & feedback in the arms dialogue. Bulletin of the Atomic Scientists, 21(1):5–10, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schurr:1965:FEA

[Sch65b] Sam H. Schurr. Forum: The economics of atomic power. Bulletin of the Atomic Scientists, 21(1):22–25, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schenck:1966:LES

[Sch66] Hilbert Schenck, Jr. Letter to the Editor: Strategy and conscience. Bulletin of the Atomic Scientists, 22(5):31–32, May 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schiller:1967:CSN

[Sch67a] Herbert I. Schiller. Communications satellites: A new institutional setting. Bulletin of the Atomic Scientists, 23(4):4–8, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schiller:1967:STV

[Sch67b] Herbert I. Schiller. The slide toward violence in the hungering world. Bulletin of the Atomic Scientists, 23(1):4–6, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SS67].

Schelling:1968:CCS

[Sch68a] Thomas C. Schelling. Current comments: Schelling responds. Bulletin of the Atomic Scientists, 24(2):28–29, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ACC^{+}67, Cro68].

Schelling:1968:FCR

[Sch68b] Thomas C. Schelling. On FAS and classified research. Bulletin of the Atomic Scientists, 24(1):63, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schiller:1968:SCI

[Sch68c] Herbert I. Schiller. Social control and individual freedom. Bulletin of the Atomic Scientists, 24(5):16–21, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schlesinger:1968:SFS

[Sch68d] James R. Schlesinger. The “Soft” factors in systems studies. Bulletin of the Atomic Scientists, 24(9):12–17, November 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneider:1968:BBM

[Sch68e] David M. Schneider. Books: The Moynihan Report and the Politics of Controversy, by Lee Rainwater and William L. Yancey. Bulletin of the Atomic Scientists, 24(3):20–21, March 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneier:1968:BBF

[Sch68f] Edward V. Schneier. Books: The Farther Shores of Politics, by George Thayer; The Great Society Reader, by Marvin Gettleman and David Mermelstein. Bulletin of the Atomic Scientists, 24 (9):36–38, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneier:1968:INP

[Sch68g] Edward V. Schneier. Intellectuals and the new politics. Bulletin of the Atomic Scientists, 24(8):15–18, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schultz:1968:WAW

[Sch68h] Theodore W. Schultz. What ails world agriculture? Bulletin of the Atomic Scientists, 24(1):28–35, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schiller:1969:BBE

[Sch69a] Herbert I. Schiller. Books: Western Economic Warfare 1947– 1967: A Case Study in Foreign Policy, by Gunnar Adler-Karlson. Bulletin of the Atomic Scientists, 25(8):44–46, October 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schurman:1969:LPC

[Sch69b] Bernard Schurman. Looking at the Peace Corps. Bulletin of the Atomic Scientists, 25(9):40–44, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1969:BMS

[Sch69c] Charles Schwartz. Berkeley: Manipulators of science — winners and losers. Bulletin of the Atomic Scientists, 25(5):21–22, 36, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1969:BRA

[Sch69d] Charles Schwartz. Book review: The atomic scientists in politics: Scientists in Politics (The Atomic Scientists Movement 1945– 56), by Donald A. Strickland. Bulletin of the Atomic Scientists, 25(6):42, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schaefer:1970:AEP

[Sch70a] Vincent J. Schaefer. Auto exhaust, pollution and weather patterns. Bulletin of the Atomic Scientists, 26(8):31–33, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schrecker:1970:BBC

[Sch70b] John Schrecker. Books: Communism and China: Ideology in Flux, by Benjamin I. Schwartz. Bulletin of the Atomic Scientists, 26(4):53, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schiller:1971:BBP

[Sch71a] Herbert I. Schiller. Books: The Pentagon Propaganda Machine, by Senator J. W. Fulbright. Bulletin of the Atomic Scientists, 27 (4):43–44, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1971:CFP

[Sch71b] Donald M. Schwartz. A comment on foreign policy thinking: Arms control and supra-nationalism. Bulletin of the Atomic Scientists, 27(4):38–41, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schriver:1972:CSH

[Sch72a] Edward Schriver. Communications: Shoreham hearings. Bulletin of the Atomic Scientists, 28(9):3, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sai72].

Schultz:1972:CR

[Sch72b] T. W. Schultz. Communications: A reply. Bulletin of the Atomic Scientists, 28(5):2, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch72d, Bet72].

Schultz:1972:ED

[Sch72c] Theodore W. Schultz. The ecosystem doom. Bulletin of the Atomic Scientists, 28(4):12–17, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [PS72].

Schultz:1972:WNE

[Sch72d] Theodore W. Schultz. Woman’s new economic commandments. Bulletin of the Atomic Scientists, 28(2):29–33, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bet72] and reply [Sch72b].

Schaeffer:1973:LC

[Sch73] Oliver A. Schaeffer. A lunar chronology. Bulletin of the Atomic Scientists, 29(10):42–45, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scheinman:1974:SNM

[Sch74a] Lawrence Scheinman. Safeguarding nuclear materials. Bulletin of the Atomic Scientists, 30(4):34–36, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schleimer:1974:DTB

[Sch74b] Joseph D. Schleimer. The day they blew up San Onofre: a scenario for sabotage at a nuclear power plant. Bulletin of the Atomic Scientists, 30(8):24–27, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mye75, Coo75b] and editorial response [Day75l].

Schwartz:1974:HPA

[Sch74c] Charles Schwartz. Helping the Pentagon aim right. Bulletin of the Atomic Scientists, 30(9):9–13, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scherer:1975:RIH

[Sch75a] Harold N. Scherer, Jr. Rebuttal: The impact of high-voltage lines. Bulletin of the Atomic Scientists, 31(7):51–52, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [YY74, You75].

Schmidt:1975:LDS

[Sch75b] Fred H. Schmidt. Letters: Defends safety study. Bulletin of the Atomic Scientists, 31(1):3–4, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day74d].

Schneider:1975:BBL

[Sch75c] Barry Schneider. Big bangs from little bombs. Bulletin of the Atomic Scientists, 31(5):24–29, May 1975. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sal75a].

Schneider:1975:SBA

[Sch75d] Barry Schneider. [Sidebar:] ‘Broken arrows’. Bulletin of the Atomic Scientists, 31(5):28, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the introduction: “A ‘broken arrow’ is the military term for an accident involving a nuclear weapon. The Pentagon has acknowledged 11 of them.”.

Schwartz:1975:CCS

[Sch75e] Charles Schwartz. The corporate connection: A surprising number of influential academic scientists are plugged into big business. Bulletin of the Atomic Scientists, 31(8):15–19, October 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mie76, Gla76d].

Schneider:1976:BSF

[Sch76a] Barry R. Schneider. Is the B-1 still flying or was it shot down this summer? Bulletin of the Atomic Scientists, 32(8):6–7, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schrock:1976:LSL

[Sch76b] Gary Schrock. Letter: A silver lining. Bulletin of the Atomic Scientists, 32(8):2, 4, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1976:LC

[Sch76c] Charles Schwartz. Letter: A challenge. Bulletin of the Atomic Scientists, 32(2):2, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneider:1977:SAC

[Sch77a] Barry R. Schneider. Stonewalling on the arms control impact statements. Bulletin of the Atomic Scientists, 33(1):5, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schultz:1977:HCC

[Sch77b] Theodore W. Schultz. The hungry, crowded competitive world. Bulletin of the Atomic Scientists, 33(8):26–31, October 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schipper:1978:BBG

[Sch78a] Les Schipper. Books: Energy: Global Prospects 1985–2000: Workshop on Alternative Energy Strategies (WAES), by Carroll Wilson. Bulletin of the Atomic Scientists, 34(3):58–60, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1978:BCN

[Sch78b] Charles Schwartz. The Berkeley controversy over nuclear weapons. Bulletin of the Atomic Scientists, 34(7):20–24, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schultz:1979:CLP

[Sch79a] Donald K. Schultz. Commentary: LNG projects. Bulletin of the Atomic Scientists, 35(4):53–54, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wol78].

Schwartz:1979:BRB

[Sch79b] David N. Schwartz. Book reviews: Arms Control and Technological Innovation, edited by Carlo Schaerf and David Carlton. Bulletin of the Atomic Scientists, 35(1):41–43, January 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1979:SDS

[Sch79c] David N. Schwartz. Salty debate in the Senate. Bulletin of the Atomic Scientists, 35(2):63–64, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schipper:1980:CCH

[Sch80a] Lee Schipper. Commentary: Conservation is here. Bulletin of the Atomic Scientists, 36(2):55–59, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schouet:1980:CWA

[Sch80b] Mauricio Schouet. Commentary: Who’s afraid of a vector? Bulletin of the Atomic Scientists, 36(6):60–62, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schroeer:1980:FRB

[Sch80c] Dietrich Schroeer. Film reviews: A is for Atom, B is for Bomb: A Portrait of Dr. Edward Teller, produced by Brian Kaufman. Bulletin of the Atomic Scientists, 36(5):55–56, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schultz:1980:EBP

[Sch80d] Theodore W. Schultz. The economics of being poor. Bulletin of the Atomic Scientists, 36(9):32–37, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1980:BRB

[Sch80e] David N. Schwartz. Book review: The Price of Defense: A New Strategy for Military Spending, by The Boston Study Group. Bulletin of the Atomic Scientists, 36(4):46–47, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneider:1981:BRB

[Sch81a] Barry Schneider. Book review: History and Memory, by Charles W. Yost. Bulletin of the Atomic Scientists, 37(6):42–44, June/ July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneider:1981:CAS

[Sch81b] Barry Schneider. Commentary: Another SALT derailment? Bulletin of the Atomic Scientists, 37(2):9, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneider:1981:CPS

[Sch81c] Barry Schneider. Commentary: Preventing star wars. Bulletin of the Atomic Scientists, 37(8):13–15, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schultz:1983:EUF

[Sch83] Theodore W. Schultz. The economics of U.S. foreign aid. Bulletin of the Atomic Scientists, 39(8):21–27, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Col84a].

Schindler:1984:CSB

[Sch84a] Saul Schindler. Commentary: On ‘Not such a bargain after all’. Bulletin of the Atomic Scientists, 40(2):64, February 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tsi83b].

Schweppe:1984:SRT

[Sch84b] John S. Schweppe. Scientific research and technology: benevolent or malevolent? Bulletin of the Atomic Scientists, 40(2):63–64, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schneider:1986:BRB

[Sch86] Stephen H. Schneider. Book review: The Machinery of Nature, by Paul R. Ehrlich. Bulletin of the Atomic Scientists, 42(9):52– 54, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schjott:1987:LFF

[Sch87] Frits Schjott. Letter: Felicitous formulation. Bulletin of the Atomic Scientists, 43(10):49, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SHT^{+}87].

Schwartz:1988:LAI

[Sch88a] Stephen Schwartz. Letter: ALPS insufficient. Bulletin of the Atomic Scientists, 44(7):62, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [CLD88].

Schweitzer:1988:WWU

[Sch88b] Glenn Schweitzer. Who wins in U.S.–Soviet science ventures? Bulletin of the Atomic Scientists, 44(8):28–32, October 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schelling:1989:ACC

[Sch89a] Thomas S. Schelling. Arms control criteria: From an airport bench. Bulletin of the Atomic Scientists, 45(4):29–31, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schoenfeld:1989:DED

[Sch89b] Gabriel Schoenfeld. A dosimeter for every dacha. Bulletin of the Atomic Scientists, 45(6):13–15, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schoenfeld:1989:TAR

[Sch89c] Gabriel Schoenfeld. Trouble aboard Red October. Bulletin of the Atomic Scientists, 45(5):13–15, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schaeffer:1990:BRS

[Sch90] Robert Schaeffer. Book review: Shades of green: Reclaiming Paradise: The Global Environmental Movement, by John Mc-Cormick. Bulletin of the Atomic Scientists, 46(3):46–47, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schaeffer:1991:LPT

[Sch91] Robert Schaeffer. Letter: Partition for “tired statesman”. Bulletin of the Atomic Scientists, 47(5):46, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Van91].

Schick:1992:LRH

[Sch92a] Paul Schick. Letter: Running hot and cold on warming. Bulletin of the Atomic Scientists, 48(7):45, September 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Leg92b].

Schoettle:1992:PUD

[Sch92b] Enid C. B. Schoettle. Perspective: U.N. dues: the price of peace. Bulletin of the Atomic Scientists, 48(5):14–16, June 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rus92b].

Schwartz:1995:NC

[Sch95a] Stephen I. Schwartz. Numbers crunch. Bulletin of the Atomic Scientists, 51(6):34, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1995:SM

[Sch95b] Stephen I. Schwartz. Sources and methods. Bulletin of the Atomic Scientists, 51(6):37, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:1996:TD

[Sch96] Stephen I. Schwartz. The $4 trillion deletion. Bulletin of the Atomic Scientists, 52(6):40, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BS97].

Schwartz:1998:AAC

[Sch98] Stephen I. Schwartz, editor. Atomic audit: the costs and consequences of U.S. nuclear weapons since 1940. Brookings Institution Press, Washington, DC, USA, 1998. ISBN 0-8157-7774-4 (hardcover), 0-8157-7773-6 (paperback). xx + 680 pp. LCCN U264.3 .A874 1998; FT AAX JTS.

Schwartz:1999:VCS

[Sch99] Stephen I. Schwartz. A very convenient scandal. Bulletin of the Atomic Scientists, 55(3):34–39, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/3/34.

Schwartz:2000:LL

[Sch00a] Stephen I. Schwartz. Lots of leaks. Bulletin of the Atomic Scientists, 56(1):67–68, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/1/67.

Schwartz:2000:MO

[Sch00b] Stephen I. Schwartz. Missed opportunities. Bulletin of the Atomic Scientists, 56(2):71–72, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/56/2/71.

Schwartz:2000:MGV

[Sch00c] Stephen I. Schwartz. Out maneuvered, out gunned, and out of view. Bulletin of the Atomic Scientists, 56(1):24–31, January/ February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
1/24.

Schwartz:2000:OOV

[Sch00d] Stephen I. Schwartz. Outmaneuvered, outgunned, and out of view. Bulletin of the Atomic Scientists, 56(1):24, January/ February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:2000:SFG

[Sch00e] Stephen I. Schwartz. Scientist, fisherman, gardener… spy? Bulletin of the Atomic Scientists, 56(6):31–38, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/6/31.

Schwartz:2001:DR

[Sch01a] Stephen I. Schwartz. Defining reliable. Bulletin of the Atomic Scientists, 57(2):55–56, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/55.

Schwartz:2001:NNC

[Sch01b] Stephen I. Schwartz. The new-nuke Chorus tunes up. Bulletin of the Atomic Scientists, 57(4):30–35, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/4/30.

Schwartz:2001:T

[Sch01c] Stephen I. Schwartz. This is not a test. Bulletin of the Atomic Scientists, 57(6):50–51, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/50.

Schwartz:2002:NYC

[Sch02a] Stephen I. Schwartz. Nukes you can use. Bulletin of the Atomic Scientists, 58(3):18–19, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/58/3/18.

Schwartz:2002:TE

[Sch02b] Stephen I. Schwartz. A tragedy of errors. Bulletin of the Atomic Scientists, 58(4):61–64, July/August 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/58/4/61.

Schwartz:2003:B

[Sch03] Stephen I. Schwartz. Bunker down. Bulletin of the Atomic Scientists, 59(3):67–70, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/3/67.

Schaper:2004:GLE

[Sch04a] Annette Schaper. Germany: Losing Europe. Bulletin of the Atomic Scientists, 60(2):32–34, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/2/32.

Schoke:2004:LER

[Sch04b] James A. Schoke. Letter to the Editor: Riding with Oppie. Bulletin of the Atomic Scientists, 60(4):4–5, July/August 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://www.highbeam.com/doc/1G1-119224463.html.

Schollmeyer:2004:PV

[Sch04c] Josh Schollmeyer. Pumping up the volume. Bulletin of the Atomic Scientists, 60(6):8–9, November/December 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schwartz:2004:SSS

[Sch04d] Stephen I. Schwartz. A slow sort of security. Bulletin of the Atomic Scientists, 60(4):??, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/18.

Schollmeyer:2005:HST

[Sch05a] Josh Schollmeyer. Homeland security: Terrorism takes a holiday. Bulletin of the Atomic Scientists, 61(4):8–9, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schollmeyer:2005:LCA

[Sch05b] Josh Schollmeyer. Lights, camera, Armageddon. Bulletin of the Atomic Scientists, 61(3):42–50, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/3/42.

Schollmeyer:2005:MR

[Sch05c] Josh Schollmeyer. Minority report. Bulletin of the Atomic Scientists, 61(1):38–39, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/1/38.

Schollmeyer:2005:SS

[Sch05d] Josh Schollmeyer. Scared straight. Bulletin of the Atomic Scientists, 61(6):70, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/6/70.

Schwartz:2005:WAF

[Sch05e] Stephen I. Schwartz. Warheads aren’t forever. Bulletin of the Atomic Scientists, 61(5):58–64, September/October 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/5/58.

Schwellenbach:2005:BPR

[Sch05f] Nick Schwellenbach. Biodefense: a plague of researchers. Bulletin of the Atomic Scientists, 61(3):14–16, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/3/14.

Schwellenbach:2005:ET

[Sch05g] Nick Schwellenbach. Empty threat. Bulletin of the Atomic Scientists, 61(5):50–57, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/5/50.

Schollmeyer:2006:CB

[Sch06a] Josh Schollmeyer. Causus bellicose. Bulletin of the Atomic Scientists, 62(2):62, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/2/62.

Schollmeyer:2006:GGS

[Sch06b] Josh Schollmeyer. Games get serious. Bulletin of the Atomic Scientists, 62(4):34–39, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/4/34.

Schollmeyer:2006:LGR

[Sch06c] Josh Schollmeyer. Lt. Gen Roméo Dallaire. Bulletin of the Atomic Scientists, 62(2):11, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schell:2007:GR

[Sch07a] Jonathan Schell. Genesis in reverse. Bulletin of the Atomic Scientists, 63(1):27–29, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/27.

Schneier:2007:MC

[Sch07b] Bruce Schneier. Master card. Bulletin of the Atomic Scientists, 63 (2):55–57, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/2/55.

Schollmeyer:2007:TC

[Sch07c] Josh Schollmeyer. The time of creation. Bulletin of the Atomic Scientists, 63(1):20, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schroeder:2007:SSS

[Sch07d] Matthew Schroeder. Sweating the small stuff. Bulletin of the Atomic Scientists, 63(5):64, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/5/64.

Scheffran:2008:CCS

[Sch08a] Jürgen Scheffran. Climate change and security. Bulletin of the Atomic Scientists, 64(2):19–25, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/2/19.

Schneider:2008:WNI

[Sch08b] Mycle Schneider. 2008 world nuclear industry status report. Bulletin of the Atomic Scientists, 64(5):26–30, November/December 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/64/5/26.

Schmid:2011:WSE

[Sch11] Sonja D. Schmid. When safe enough is not good enough: Organizing safety at Chernobyl. Bulletin of the Atomic Scientists, 67 (2):19–29, March/April 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/2/19.

Schreurs:2012:PP

[Sch12] Miranda A. Schreurs. The politics of phase-out. Bulletin of the Atomic Scientists, 68(6):30–41, November/December 2012. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/68/6/30.

Schneider:2013:FGE

[Sch13a] Mycle Schneider. France’s great energy debate. Bulletin of the Atomic Scientists, 69(1):27–35, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/69/1/27.

Schneider:2013:NPF

[Sch13b] Mycle Schneider. Nuclear power and the French energy transition: It’s the economics, stupid! Bulletin of the Atomic Scientists, 69(1):18–26, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/69/1/18.

Schlosser:2015:TND

[Sch15a] Eric Schlosser. Today’s nuclear dilemma. Bulletin of the Atomic Scientists, 71(6):11–17, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schmidt:2015:WSC

[Sch15b] Gavin A. Schmidt. What should climate scientists advocate for? Bulletin of the Atomic Scientists, 71(1):70–74, January 2015. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schmid:2016:WIT

[Sch16] Sonja D. Schmid. What if there’s a next time? Preparedness after Chernobyl and Fukushima. Bulletin of the Atomic Scientists, 72 (4):260–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scharre:2018:HSW

[Sch18] Paul Scharre. How swarming will change warfare. Bulletin of the Atomic Scientists, 74(6):385–389, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1533209.

Scheffran:2019:ECW

[Sch19] Jürgen Scheffran. The entwined Cold War roots of missile defense and climate geoengineering. Bulletin of the Atomic Scientists, 75(5):222–228, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1654256.

Scheirer:2020:PBS

[Sch20] Walter Scheirer. A pandemic of bad science. Bulletin of the Atomic Scientists, 76(4):175–184, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1778361.

Schoonover:2021:CCS

[Sch21] Rod Schoonover. Climate change should be recognized for what it is: an issue of national security. Bulletin of the Atomic Scientists, 77(1):31–32, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1860330.

Scheirer:2024:AMD

[Sch24] Walter J. Scheirer. AI misinformation detectors can’t save us from tyranny — at least not yet. Bulletin of the Atomic Scientists, 80(5):308–313, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2388467.

Strauch:2020:HPU

[SCHD20] Yonatan Strauch, Angela Carter, and Thomas Homer-Dixon. However the pandemic unfolds, it’s time for oil use to peak-and society to prepare for the fallout. Bulletin of the Atomic Scientists, 76(5):238–243, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1806577.

STC:1946:STA

[Sci46] Scientific and Technical Committee of the UN Atomic Energy Commission. The scientific and technical aspects of atomic energy control … a first report. Bulletin of the Atomic Scientists, 2(7– 8):6–13, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SCLP:1948:FLC

[Sci48] Scientists’ Committee on Loyalty Problems. FAS Loyalty Committee summarizes clearance procedures. Bulletin of the Atomic Scientists, 4(12):384, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SLSGWASFAS:1949:UPN

[Sci49] Science Legislation Study Group, Washington Association of Scientists, FAS. Uncertain prospect: National Science Foundation. Bulletin of the Atomic Scientists, 5(1):25–26, 32, January 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SMACNSRB:1951:TRS

[Sci51] Scientific Manpower Advisory Committee of the National Security Resources Board. The Thomas report on scientific manpower. Bulletin of the Atomic Scientists, 7(3):89–93, March 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SCLS:1955:LUP

[Sci55a] Scientists’ Committee on Loyalty and Security. Loyalty and the U.S. Public Health Service Grants. Bulletin of the Atomic Scientists, 11(5):196–197, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SCLS:1955:SIC

[Sci55b] Scientists’ Committee on Loyalty and Security. Some individual cases. Bulletin of the Atomic Scientists, 11(4):151–155, 158, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

SEPP:1992:DW

[Sci92] Science and Environmental Policy Project. Dissent on warming. Bulletin of the Atomic Scientists, 48(5):37, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sclove:1979:BRB

[Scl79] Richard Sclove. Book reviews: Risk Assessment Review Group Report to the U.S. Nuclear Regulatory Commission, by H. W. Lewis and others. Bulletin of the Atomic Scientists, 35(2):46– 48, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sclove:1982:DMD

[Scl82] Richard Sclove. Decision-making in a democracy. Bulletin of the Atomic Scientists, 38(5):44–49, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gib82].

Scott:1966:LES

[Sco66] J. P. Scott. Letter to the Editor: Sociobiology. Bulletin of the Atomic Scientists, 22(2):33–34, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Guh65].

Scoville:1970:VNA

[Sco70] Herbert Scoville, Jr. Verification of nuclear arms limitations: An analysis. Bulletin of the Atomic Scientists, 26(8):6–11, October 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1971:BBM

[Sco71] Herbert Scoville, Jr. Books: The Military Establishment: Its Impact on American Society, by Adam Yarmolinsky. Bulletin of the Atomic Scientists, 27(6):38, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1972:IAL

[Sco72] Herbert Scoville, Jr. Issues in arms limitations — four views: I. Next steps in limiting strategic arms. Bulletin of the Atomic Scientists, 28(3):9–12, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1974:BBA

[Sco74] Herbert Scoville, Jr. Books: American Arms and a Changing Europe: Dilemmas of Deterrence and Disarmament, by Warner R. Schilling, William T. R. Fox, Catherine M. Kelleher, and Donald J. Puchala. Bulletin of the Atomic Scientists, 30(9):47–48, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1977:SAR

[Sco77] Herbert Scoville, Jr. Slowing the arms race. Bulletin of the Atomic Scientists, 33(7):4–5, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1979:BRB

[Sco79] Herbert Scoville, Jr. Book review: Defense Politics: a Budgetary Perspective, by Arnold Kanter. Bulletin of the Atomic Scientists, 35(7):44, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scott:1984:NWG

[Sco84a] Robert T. Scott. Now a “warhead” gap. Bulletin of the Atomic Scientists, 40(9):43–44, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1984:PCC

[Sco84b] Herbert Scoville, Jr. Perspective: Congressional cliffhanger. Bulletin of the Atomic Scientists, 40(8):5–6, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scott:1989:LUM

[Sco89] Douglas Scott. Letter: U.N. for multilateral treaties. Bulletin of the Atomic Scientists, 45(7):45, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hei89].

Scott:1990:LRS

[Sco90] Douglas Scott. Letter: Reactor sale does not break treaty. Bulletin of the Atomic Scientists, 46(8):46, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gol90a].

Scott:1992:LBA

[Sco92] Terry Scott. Letter: Bulletin abandoning human rights? Bulletin of the Atomic Scientists, 48(8):47, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rad92a].

Scott:1994:LRC

[Sco94a] Terry Scott. Letter: Reality check. Bulletin of the Atomic Scientists, 50(3):60, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scott:1994:LSM

[Sco94b] Terry Scott. Letter: Selective memory. Bulletin of the Atomic Scientists, 50(5):3, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gey94].

Scott:1995:LTN

[Sco95] Terry Scott. Letter: Thanks but no thanks. Bulletin of the Atomic Scientists, 51(2):71, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fos95].

Scoblic:2008:DR

[Sco08a] J. Peter Scoblic. Disarmament redux. Bulletin of the Atomic Scientists, 64(1):34–39, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/1/34.

Scoblic:2008:DRU

[Sco08b] J. Peter Scoblic. Disarmament redux: The U.S. foreign policy establishment is beginning to consider progress toward ‘the d-word’ above and beyond deterrence a global security imperative. Bulletin of the Atomic Scientists, 64(1):34–57, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scoville:1979:C

[SCT^{+}79] Herbert Scoville, John C. Culver, Bardyl R. Tirana, Bernard T. Feld, Henry Kendall, Richard Barnet, John McDonough, Theodore A. Postol, Sanford Gottlieb, Kay Camp, A. S. Collins, Jr., John Conner, Annette LaRocque, George Kistiakowsky, and I. F. Stone. Comments. Bulletin of the Atomic Scientists, 35 (7):40–43, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Scudder:1965:KCM

[Scu65] Thayer Scudder. The Kariba case: Manmade lakes and resource development in Africa. Bulletin of the Atomic Scientists, 21(10): 6–11, December 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Slone:1955:SMH

[SD55] Ralph W. Slone and Wilbur G. Dunning. The scientific method in human affairs: A reply to Niebuhr’s “Limitations of the Scientific Method”; a reply to Auger’s article. Bulletin of the Atomic Scientists, 11(6):227–229, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nie55, Aug55] and comment [Kir55].

Socolar:1956:LEP

[SD56] Sidney J. Socolar and William C. Davidon. Letter to the Editor: Proposals for cessation of H-bomb testing. Bulletin of the Atomic Scientists, 12(9):353, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simon:1983:CDB

[SD83] Julian L. Simon and Herman E. Daly. Commentary: In defense: The Ultimate Resource. Bulletin of the Atomic Scientists, 39(5): 57–59, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dal82] and comment [Har83].

Schwartz:1986:ACM

[SD86] William A. Schwartz and Charles Derber. Arms control: misplaced focus. Bulletin of the Atomic Scientists, 42(3):39–44, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stremmel:1987:LMS

[SD87] William J. Stremmel and Lloyd J. Dumas. Letter: Military spending debate. Bulletin of the Atomic Scientists, 43(1):57– 58, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dum86].

Schlesinger:1988:GG

[SD88] James Schlesinger and Michael Dukakis. GWEN and the governor. Bulletin of the Atomic Scientists, 44(9):23, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:2009:HBC

[SD09] Peter A. Singer and Abdallah S. Daar. How biodevelopment can enhance biosecurity. Bulletin of the Atomic Scientists, 65(2): 23–30, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/65/2/23.

Smith:2010:ASB

[SD10a] Geoffrey L. Smith and Neil Davison. Assessing the spectrum of biological risks. Bulletin of the Atomic Scientists, 66(1):1–11, January 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
1/1.

Stockton:2010:DDW

[SD10b] Peter Stockton and Ingrid Drake. From danger to dollars: What the US should do with its highly enriched uranium. Bulletin of the Atomic Scientists, 66(6):43–55, November/December 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/6/43.

Smetana:2015:MMT

[SD15] Michal Smetana and Ondrej Ditrych. The more the merrier: Time for a multilateral turn in nuclear disarmament. Bulletin of the Atomic Scientists, 71(3):30–37, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schmidt:2017:SAT

[SD17] Gavin A. Schmidt and Simon D. Donner. Scientific advocacy: A tool for assessing the risks of engagement. Bulletin of the Atomic Scientists, 73(5):344–347, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1364008.

Shizume:1962:LEJ

[SdSP62] Eri Yagi Shizume and Derek J. de Solla Price. Letter to the Editor: Japanese bomb. Bulletin of the Atomic Scientists, 18(9): 29, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Selove:1958:IRM

[SE58] Walter Selove and Mortimer M. Elkind. Introduction: Radiation and man. Bulletin of the Atomic Scientists, 14(1):7–8, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sears:1956:LEM

[Sea56] M. J. Sears. Letter to the Editor: Mohawk Association of Scientists and Engineers. Bulletin of the Atomic Scientists, 12(10): 381, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sears:1961:PT

[Sea61] Paul B. Sears. The perspective of time. Bulletin of the Atomic Scientists, 17(8):322–325, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1963:CG

[Sea63a] Glenn T. Seaborg. A common ground. Bulletin of the Atomic Scientists, 19(10):42–44, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1963:ICV

[Sea63b] Glenn T. Seaborg. International cooperation: The value of exchange. Bulletin of the Atomic Scientists, 19(7):25, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1965:UEC

[Sea65] Glenn T. Seaborg. The U.S. [energy] capacity — in 1980. Bulletin of the Atomic Scientists, 21(5):24, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1967:WAI

[Sea67] Glenn T. Seaborg. What’s ahead for international science? Bulletin of the Atomic Scientists, 23(1):24–28, January 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1968:NWF

[Sea68] Glenn T. Seaborg. Need we fear our nuclear future? Bulletin of the Atomic Scientists, 24(1):36–42, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seabury:1969:AK

[Sea69] Paul Seabury. Into the arms of Krupp. Bulletin of the Atomic Scientists, 25(9):39–40, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seabord:1970:IC

[Sea70a] Glenn T. Seabord. An international challenge. Bulletin of the Atomic Scientists, 26(9):5–7, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1970:PPR

[Sea70b] Glenn T. Seaborg. Part 1: Projection and recollection: Our nuclear future — 1995. Bulletin of the Atomic Scientists, 26(6): 7–14, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1971:MA

[Sea71] Glenn T. Seaborg. On misunderstanding the atom. Bulletin of the Atomic Scientists, 27(7):46–53, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gad72].

Seaborg:1972:VS

[Sea72a] Glenn T. Seaborg. Vikram A. Sarabhai. Bulletin of the Atomic Scientists, 28(4):48, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sears:1972:BBP

[Sea72b] William R. Sears. Books: Project Paperclip, by Clarence G. Lasby. Bulletin of the Atomic Scientists, 28(6):55–56, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1973:ANW

[Sea73] Glenn T. Seaborg. In the Americas: A new world outlook. Bulletin of the Atomic Scientists, 29(8):3, 22–27, October 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1979:AER

[Sea79] Glenn T. Seaborg. Albert Einstein – a reflection. Bulletin of the Atomic Scientists, 35(3):20, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Einstein and peace.

Seaborg:1985:PAB

[Sea85a] Glenn T. Seaborg. Premonitions after the bombs. Bulletin of the Atomic Scientists, 41(11):31–33, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sears:1985:LMS

[Sea85b] John Sears. Letter: Military spending and economic pumppriming. Bulletin of the Atomic Scientists, 41(2):56, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bal84b].

Seaborg:1989:ACL

[Sea89a] Glenn T. Seaborg. Arms control at the labs. Bulletin of the Atomic Scientists, 45(6):11, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seaborg:1989:PWL

[Sea89b] Glenn T. Seaborg. Perspective: Weapons labs need new thinking. Bulletin of the Atomic Scientists, 45(6):10–12, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sea:2001:RR

[Sea01] Geoffrey Sea. Research in the rainforest. Bulletin of the Atomic Scientists, 57(1):71–73, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/1/71.

Seelman:1979:EPN

[See79] Katherine D. Seelman. Energy policy of The National Council of Churches. Bulletin of the Atomic Scientists, 35(8):10–11, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Segre:1970:EFPa

[Seg70a] Emilio Segrè. Enrico Fermi: Physicist. Bulletin of the Atomic Scientists, 26(9):32, 37–39, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from [Seg70b].

Segre:1970:EFPb

[Seg70b] Emilio Segrè. Enrico Fermi: physicist. University of Chicago Press, Chicago, IL, USA and London, UK, 1970. ISBN 0-226-74472-8. x + 276 pp. LCCN QC16.F46 S4.

Segal:2013:CTC

[Seg13] Adam Segal. The code not taken: China, the United States, and the future of cyber espionage. Bulletin of the Atomic Scientists, 69(5):38–45, September 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/5/38.

Seitz:1949:AER

[Sei49] Frederick Seitz. The atomic explosion in Russia: The danger ahead. Bulletin of the Atomic Scientists, 5(10):266, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1950:PCW

[Sei50] Frederick Seitz. Physicists and the Cold War. Bulletin of the Atomic Scientists, 6(3):83–89, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1953:ODW

[Sei53] Frederick Seitz. Offensive or defensive weapons? Bulletin of the Atomic Scientists, 9(9):325–327, 336, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1960:BBA

[Sei60] Frederick Seitz. Books: NATO in the 1960’s, by Alastair Buchan; NATO and American Security, edited by Klaus Knorr. Bulletin of the Atomic Scientists, 16(9):370–371, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1962:BBB

[Sei62] Frederick Seitz. Books: The Balance of Terror, by Pierre Gallois, with a foreword by Raymond Aron. Bulletin of the Atomic Scientists, 18(4):41, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seith:1969:BBW

[Sei69] Alex R. Seith. Books: Chemical Warfare: A Study in Restraints, by Frederic J. Brown. Bulletin of the Atomic Scientists, 25(5): 52–53, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seidman:1971:BTI

[Sei71a] Aaron Seidman. Barriers to technical innovation. Bulletin of the Atomic Scientists, 27(3):29–31, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seith:1971:BBL

[Sei71b] Alex R. Seith. Books: Legal Limits on the Use of Chemical and Biological Weapons, by Ann Van Wynen Thomas and A.

J. Thomas, Jr. Bulletin of the Atomic Scientists, 27(4):42–43, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seith:1973:BBC

[Sei73] Alex R. Seith. Books: The Concept of Aggression in International Law, by Ann Van Wynen Thomas and A. J. Thomas. Bulletin of the Atomic Scientists, 29(9):56, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1980:RU

[Sei80] Frederick Seitz. Reflections on UNCSTD. Bulletin of the Atomic Scientists, 36(5):48–51, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seitz:1981:BRB

[Sei81] Frederick Seitz. Book review: America’s Technology Slip, by Simon Ramo. Bulletin of the Atomic Scientists, 37(5):43–44, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seiler:1984:CCL

[Sei84] Ralph Seiler. Commentary: Chicken Little. Bulletin of the Atomic Scientists, 40(5):64, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seldon:1959:LES

[Sel59] Percy Seldon. Letter to the Editor: Secret biological warfare possible? Bulletin of the Atomic Scientists, 15(10):401, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seligman:1963:BBR

[Sel63] Eustace Seligman. Books: Defense and Retaliation: A German View, by Helmut Schmidt. Bulletin of the Atomic Scientists, 19(3):25–26, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Selk:1969:FPN

[Sel69a] Merry Selk. The five principles — a new approach. Bulletin of the Atomic Scientists, 25(2):78, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Selk:1969:SBT

[Sel69b] Merry Selk. Sentinel in the backyard: The transitional reaction. Bulletin of the Atomic Scientists, 25(1):7, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Selk:1969:SHS

[Sel69c] Merry Selk. Styles of handling student demonstrations. Bulletin of the Atomic Scientists, 25(6):36–39, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Selbin:1977:RUT

[Sel77] Joel Selbin. Rebuttal: Unreal thinking about energy. Bulletin of the Atomic Scientists, 33(7):54–55, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet77b, Bet77a].

Sellevaag:2000:NMD

[Sel00] Inge Sellevåg. National missile defense: Vardø exposed. Bulletin of the Atomic Scientists, 56(2):26–29, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/2/26.

Sellstrom:2021:USE

[Sel21] Åke Sellström. UNSCOM: A successful experiment in disarmament. Bulletin of the Atomic Scientists, 77(4):177–179, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1941543.

Semyonov:1953:ESI

[Sem53] Nikolai Semyonov. The exchange of scientific information with the Soviet Union. Bulletin of the Atomic Scientists, 9(1):13–14, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Semenov:1959:FMA

[Sem59] N. N. Semenov. The future of man in the Atomic Age. Bulletin of the Atomic Scientists, 15(3):123–126, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Semenov:1964:WF

[Sem64] N. N. Semenov. The world of the future. Bulletin of the Atomic Scientists, 20(2):10–15, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sagan:1970:CCP

[SEM70] Leonard A. Sagan, Robert B. Egbert, and R. C. Mallatt. Communications: Cars and pollution. Bulletin of the Atomic Scientists, 26(5):43–44, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mil71].

McMahon:1947:MSS

[Sen47] Senator McMahon. McMahon suggestions solution of veto controversy through World Court. Bulletin of the Atomic Scientists, 3(2):48–49, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From a January 27 1947 speech to the U.S. Senate.

SFRSD:1957:RDU

[Sen57] Senate Foreign Relations Subcommittee on Disarmament. The role of disarmament in United States foreign policy. Bulletin of the Atomic Scientists, 13(4):139–142, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Senter:1963:DM

[Sen63] Raymond D. Senter. The dilemma of the military. Bulletin of the Atomic Scientists, 19(10):24–27, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sen:1965:LDP

[Sen65a] Dipak K. Sen. Letters: Disarmament proposal. Bulletin of the Atomic Scientists, 21(2):35–36, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Senn:1965:LEF

[Sen65b] Édouard Senn. Letter to the Editor: From a French friend. Bulletin of the Atomic Scientists, 21(6):49–51, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Senn:1967:CFV

[Sen67] Édouard Senn. Comments: A French view of Vietnam. Bulletin of the Atomic Scientists, 23(2):26–27, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Senghaas:1987:CFE

[Sen87] Dieter Senghaas. Conventional forces in Europe: Dismantle offense, strengthen defense. Bulletin of the Atomic Scientists, 43 (10):9–11, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sepasspour:2023:RCW

[Sep23] Rumtin Sepasspour. A reality check and a way forward for the global governance of artificial intelligence. Bulletin of the Atomic Scientists, 79(5):304–315, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2245249.

Sessler:1983:BRB

[Ses83] Andrew M. Sessler. Book review: Refusnik: Trapped in the Soviet Union, by Mark Ya Azbel. Bulletin of the Atomic Scientists, 39(7):44–45, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sethi:2016:WIT

[Set16] Manpreet Sethi. What if there’s a next time? Preparedness after Chernobyl and Fukushima. Bulletin of the Atomic Scientists, 72 (4):262–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sewell:1967:PPP

[Sew67] W. R. Derrick Sewell. Pipedream or practical possibility? [Nawapa]. Bulletin of the Atomic Scientists, 23(7):9–13, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seznec:2016:SAS

[Sez16] Jean-François Seznec. Saudi Arabia’s sell-off of Aramco: Risk or opportunity? Bulletin of the Atomic Scientists, 72(6):378– 383, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seznec:2020:WUS

[Sez20] Jean-Francois Seznec. Why US–Saudi Arabia relations will continue to be close, even when climate action reduces demand for oil. Bulletin of the Atomic Scientists, 76(5):244–248, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1806578.

Sheldon:1977:LT

[SF77] John Sheldon and Bernard T. Feld. Letter: On technocrats. Bulletin of the Atomic Scientists, 33(5):2, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel77a].

Schware:1981:CDH

[SF81a] Robert Schware and Edward J. Friedman. Climate debate heats up. Bulletin of the Atomic Scientists, 37(10):31–33, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The article includes a figure showing an engraving of the Argentière glacier in the French Alps in the 1850s (the ‘Little Ice Age’) and a 1966 photograph of the glacier showing its dramatic retreat.

Sommerville:1981:CCS

[SF81b] John Sommerville and Bernard T. Feld. Commentary: A credible Soviet proposal. Bulletin of the Atomic Scientists, 37(10): 57, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shrader-Frechette:1994:HLW

[SF94] Kristin Shrader-Frechette. High-level waste, low-level logic. Bulletin of the Atomic Scientists, 50(6):40–45, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CSF95, CBSF95].

Schneider:2012:WNI

[SF12] Mycle Schneider and Antony Froggatt. 2011–2012 world nuclear industry status report. Bulletin of the Atomic Scientists, 68(5):8– 22, September/October 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/5/8.

Schneider:2014:WNI

[SF14] Mycle Schneider and Antony Froggatt. 2012–2013 world nuclear industry status report. Bulletin of the Atomic Scientists, 70(1):70–84, January 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/70/1/70.full.pdf+html.

Schneider:2011:WNI

[SFT11] Mycle Schneider, Antony Froggatt, and Steve Thomas. 2010– 2011 world nuclear industry status report. Bulletin of the Atomic Scientists, 67(4):60–77, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/67/4/60.

Szent-Gyorgyi:1959:LEP

[SG59] Albert Szent-Györgyi. Letter to the Editor: The price of peace. Bulletin of the Atomic Scientists, 15(8):346–347, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szent-Gyorgyi:1962:LES

[SG62a] Albert Szent-Györgyi. Letter to the Editor: Solving the unsolvable. Bulletin of the Atomic Scientists, 18(8):32–33, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szent-Gyorgyi:1962:LEU

[SG62b] Albert Szent-Györgyi. Letter to the Editor: Ultimate goal of disarmament. Bulletin of the Atomic Scientists, 18(6):43, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szent-Gyorgyi:1964:BMP

[SG64] Albert Szent-Györgyi. The brain, morals, and politics. Bulletin of the Atomic Scientists, 20(5):2–3, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szent-Gyorgyi:1965:LEN

[SG65] Albert Szent-Györgyi. Letter to the Editor: Nursery rhyme. Bulletin of the Atomic Scientists, 21(3):44–45, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stern:1969:OCS

[SG69a] Philip M. Stern and Harold P. Green. The Oppenheimer case: security on trial. Harper & Row, New York, NY, USA, 1969. xii + 591 pp. LCCN QC16.O62 S69. With a special commentary by Lloyd K. Garrison.

Szent-Gyorgyi:1969:DWM

[SG69b] Albert Szent-Györgyi. The dual world of man: Reflections on science and government. Bulletin of the Atomic Scientists, 25 (8):33–34, 37, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szent-Gyorgyi:1969:SBC

[SG69c] Albert Szent-Györgyi. Science and budget cutting. Bulletin of the Atomic Scientists, 25(10):16–17, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stern:1971:OCS

[SG71] Philip M. Stern and Harold P. Green. The Oppenheimer case: security on trial. Hart-Davis, London, UK, 1971. ISBN 0-246-64035-9. xiv + 591 pp. LCCN QC16.O62 S69 1971. Foreword by Sir Solly Zuckerman.

Szent-Gyorgyi:1975:ELC

[SG75] Albert Szent-Györgyi. Editorial: A little catechism. Bulletin of the Atomic Scientists, 31(4):5, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [MD75].

Schindler:1983:CWP

[SG83] Saul Schindler and Richard L. Garwin. Commentary: Who proposes, who disposes, who pays? Bulletin of the Atomic Scientists, 39(7):62–64, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gar83b].

Strausz-Hupe:1961:LER

[SH61] Robert Strausz-Hupé. Letter to the Editor: For the record. Bulletin of the Atomic Scientists, 17(10):428–429, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LM61b].

Sawhill:1989:LWA

[SH89a] John Sawhill and Kathleen Hart. Letters: Wrong about Whitehead. Bulletin of the Atomic Scientists, 45(5):44–45, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Har89b].

Stern:1989:LSS

[SH89b] Paul C. Stern and Jo L. Husbands. Liberating Soviet social science. Bulletin of the Atomic Scientists, 45(7):28–31, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pow90c].

Shurcliff:1994:LLF

[SH94] William A. Shurcliff and Thomas A. Halsted. Letters: Latest fallout from Operation Crossroads. Bulletin of the Atomic Scientists, 50(4):3, 58, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schaper:1995:LTV

[SH95] Annette Schaper and Mark Hibbs. Letters: Two views. Bulletin of the Atomic Scientists, 51(3):69, May/June 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hib95d].

Shapiro:1950:BRA

[Sha50] Maurice M. Shapiro. Book review: Atomic Energy, edited by R. E. Peierls and J. L. Crammer. Bulletin of the Atomic Scientists, 6(11):347, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shaffer:1954:CC

[Sha54a] Aaron Shaffer. Current comment. Bulletin of the Atomic Scientists, 10(5):171, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1954:BBG

[Sha54b] Harlow Shapley. Books: Galileo Galilei: Dialogue Concerning the Two Chief World Systems — Ptolemaic and Copernican, translated by Stillman Drake with a foreword by Albert Einstein; Galileo Galilei: Dialogue on the Great World Systems, revised by Giorgio de Santillana. Bulletin of the Atomic Scientists, 10 (5):162, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1959:BBM

[Sha59] Malcolm Sharp. Books: No More War, by Linus Pauling; World Peace Through World Law, by Grenville Clark and Louis B. Sohn. Bulletin of the Atomic Scientists, 15(1):44–46, January 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1960:AD

[Sha60] Malcolm P. Sharp. Agreement and disarmament. Bulletin of the Atomic Scientists, 16(5):179–180, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1961:GUD

[Sha61] Malcolm P. Sharp. Graduated unilateral disarmament. Bulletin of the Atomic Scientists, 17(3):113–114, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Osg60].

Sharlin:1963:SB

[Sha63] Harold I. Sharlin. The scientist in biography. Bulletin of the Atomic Scientists, 19(9):27–28, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shafroth:1965:LED

[Sha65] Steven Shafroth. Letter to the Editor: In defense of MLF. Bulletin of the Atomic Scientists, 21(3):44, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapiro:1966:LEC

[Sha66] Herbert Shapiro. Letter to the Editor: On China. Bulletin of the Atomic Scientists, 22(1):37–38, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharma:1969:PSN

[Sha69] Ravi Dutta Sharma. Possibility of successful nuplex for India’s Thar desert. Bulletin of the Atomic Scientists, 25(9):31, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1973:CS

[Sha73] Malcolm Sharp. Concessions for survival. Bulletin of the Atomic Scientists, 29(9):48–49, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1978:IRUa

[Sha78a] Jane M. O. Sharp. Isaiah revisited: U.N. Special Session on Disarmament: I: The case for national initiatives. Bulletin of the Atomic Scientists, 34(4):18–21, 52–59, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1978:IRUb

[Sha78b] Jane M. O. Sharp. Isaiah revisited: U.N. Special Session on Disarmament. Part II. Feasible initiatives for the United States. Bulletin of the Atomic Scientists, 34(6):24–31, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1980:CSI

[Sha80] Deborah Shapley. Commentary: SALT is down, but is it out? Bulletin of the Atomic Scientists, 36(3):6–7, March 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1981:BRB

[Sha81a] Jane Sharp. Book review: National Disarmament Mechanisms, by Larry M. Ross and John R. Redick. Bulletin of the Atomic Scientists, 37(4):46–47, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [RR81b].

Sharp:1981:ACP

[Sha81b] Jane M. O. Sharp. Arms control priorities. Bulletin of the Atomic Scientists, 37(7):26–30, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1981:PGA

[Sha81c] Malcolm P. Sharp. Perspective: Gambling on amity. Bulletin of the Atomic Scientists, 37(5):1, 4, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1982:NWA

[Sha82] Jane M. O. Sharp. Nuclear weapons and Alliance cohesion. Bulletin of the Atomic Scientists, 38(6):33–36, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1983:P

[Sha83] Jane M. O. Sharp. A perspective. Bulletin of the Atomic Scientists, 39(10):12–13, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shanebrook:1984:CTM

[Sha84a] J. Richard Shanebrook. Commentary: On ‘Two missiles in every garage’. Bulletin of the Atomic Scientists, 40(1):61, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste83b].

Shapley:1984:BRB

[Sha84b] Deborah Shapley. Book review: Nuclear War, Nuclear Peace, by Leon Wieseltier. Bulletin of the Atomic Scientists, 40(5):47–48, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1984:PA

[Sha84c] Deborah Shapley. Pax Antarctica. Bulletin of the Atomic Scientists, 40(6):30–33, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1984:WWV

[Sha84d] Deborah Shapley. What would Vannevar Bush say? Bulletin of the Atomic Scientists, 40(3):19–20, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cha84].

Sharp:1984:PSR

[Sha84e] Jane M. O. Sharp. Perspective: Soviet response to cruise and Pershing. Bulletin of the Atomic Scientists, 40(3):3–4, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shanebrook:1985:LNI

[Sha85a] J. Richard Shanebrook. Letter: Nuclear intimidation. Bulletin of the Atomic Scientists, 41(2):57, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ger84].

Sharp:1985:LWP

[Sha85b] Jane M. O. Sharp. Letter: Warsaw Pact. Bulletin of the Atomic Scientists, 41(6):55–56, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sta85].

Sharp:1985:RNN

[Sha85c] Jane M. O. Sharp. Reshaping NATO nuclear policy. Bulletin of the Atomic Scientists, 41(4):38–44, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hat85].

Sharlet:1986:SLD

[Sha86] Robert Sharlet. Soviet Law and the Daniloff case. Bulletin of the Atomic Scientists, 42(10):14–18, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lar87].

Shallhorn:1987:CDS

[Sha87a] Steve Shallhorn. Canadian defense: Standing up to the United States. Bulletin of the Atomic Scientists, 43(8):16–17, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shanebrook:1987:ENW

[Sha87b] J. Richard Shanebrook. Endless nuclear weapons. Bulletin of the Atomic Scientists, 43(2):63–64, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tay86].

Shapley:1987:JWA

[Sha87c] Deborah Shapley. Japan’s wartime atomic research. Bulletin of the Atomic Scientists, 43(9):8, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1987:IVN

[Sha87d] Jane M. O. Sharp. International views: NATO’s security dilemma. Bulletin of the Atomic Scientists, 43(2):42–44, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shalat:1990:BRE

[Sha90] Stuart L. Shalat. Book review: ELFs get shock treatment: Currents of Death: Power Lines, Computer Terminals, and the Attempt to Cover Up Their Threat to Your Health, by Paul Brodeur. Bulletin of the Atomic Scientists, 46(3):47–48, April 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jau90b].

Shapley:1991:BRL

[Sha91a] Deborah Shapley. Book review: Like Ike: The Commanders, by Bob Woodward. Bulletin of the Atomic Scientists, 47(7):41– 42, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1991:BRM

[Sha91b] Deborah Shapley. Book review: Mr. Cold War: As I Saw It, by Dean Rusk, as told to Richard Rusk and Daniel S. Papp. Bulletin of the Atomic Scientists, 47(6):38–39, July/August 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharma:1991:ILS

[Sha91c] Dhirendra Sharma. India’s lopsided science. Bulletin of the Atomic Scientists, 47(4):32–36, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapiro:1992:LDN

[Sha92a] Maurice M. Shapiro. Letter: Duke next? Bulletin of the Atomic Scientists, 48(2):48, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ade91].

Sharp:1992:INW

[Sha92b] Jane Sharp. If not NATO, who? Bulletin of the Atomic Scientists, 48(8):29–32, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1992:VB

[Sha92c] Jane Sharp. A view from Britain. Bulletin of the Atomic Scientists, 48(4):20–21, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1993:CSP

[Sha93a] Deborah Shapley. Clintonizing science policy. Bulletin of the Atomic Scientists, 49(10):39–43, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1993:END

[Sha93b] Jane Sharp. Europe’s nuclear dominos. Bulletin of the Atomic Scientists, 49(5):29–33, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jon93b].

Sharp:1993:PWM

[Sha93c] Jane Sharp. Perspective: The West’s moral failure. Bulletin of the Atomic Scientists, 49(2):11, 47, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dri93].

Sharp:1994:PBB

[Sha94] Jane M. O. Sharp. Perspective: The Bill and Boris Show. Bulletin of the Atomic Scientists, 50(1):12–13, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shapley:1995:MIW

[Sha95a] Deborah Shapley. McNamara’s inner war. Bulletin of the Atomic Scientists, 51(4):13–15, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1995:BRM

[Sha95b] Jane M. O. Sharp. Book review: Missed opportunities: Ending Europe’s Wars, by Jonathan Dean. Bulletin of the Atomic Scientists, 51(1):61–63, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1995:PLM

[Sha95c] Jane M. O. Sharp. Perspective: Let’s make a deal: NATO and CFE. Bulletin of the Atomic Scientists, 51(2):19–21, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1996:GOB

[Sha96] Jane M. O. Sharp. Guest opinion: Bosnia needs a friend with staying power. Bulletin of the Atomic Scientists, 52(2):4, March/ April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1997:PBB

[Sha97] Jane M. O. Sharp. Perspective: Bosnia: Begin again. Bulletin of the Atomic Scientists, 53(2):17–19, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharma:1998:HB

[Sha98a] Kalpana Sharma. The Hindu bomb. Bulletin of the Atomic Scientists, 54(4):30–33, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1998:BRB

[Sha98b] Jane M. O. Sharp. Book review: To End a War, by Richard Holbrooke. Bulletin of the Atomic Scientists, 54(5):62–63, September/October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharp:1998:NSS

[Sha98c] Jane M. O. Sharp. NATO: Spreading the security blanket. Bulletin of the Atomic Scientists, 54(1):23–27, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharma:1999:KWR

[Sha99] Kalpana Sharma. Kashmir: Wilting rose. Bulletin of the Atomic Scientists, 55(1):4, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/4.

Shaw:2000:TIB

[Sha00] Jeff Shaw. Trident II: $9 billion for what? Bulletin of the Atomic Scientists, 56(4):13–14, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/4/13.

Shambroom:2006:FS

[Sha06] Paul Shambroom. The faces of security. Bulletin of the Atomic Scientists, 62(5):46–53, September/October 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/5/46.

Shaw:2010:LRH

[Sha10] Douglas B. Shaw. Lessons of restraint: How Canada helps explain and strengthen the nonproliferation norm. Bulletin of the Atomic Scientists, 66(5):45–54, September/October 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/5/45.

Sharma:2014:IRH

[Sha14] Kalpana Sharma. The Indian response: How to energize women. Bulletin of the Atomic Scientists, 70(2):13–16, March 2014. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sharikov:2018:AIC

[Sha18] Pavel Sharikov. Artificial intelligence, cyberattack, and nuclear weapons — a dangerous combination. Bulletin of the Atomic Scientists, 74(6):368–373, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2018.1533185.

Sharma:2022:CMD

[Sha22] Ishan Sharma. Creating a model democratic alternative to the surveillance state. Bulletin of the Atomic Scientists, 78(5):249– 255, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2109325.

Sherwin:1956:SPT

[She56] C. W. Sherwin. Securing peace through military technology. Bulletin of the Atomic Scientists, 12(5):159–163, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherwin:1960:BBM

[She60] C. W. Sherwin. Books: Strategy in the Missile Age, by Bernard Brodie; The Question of National Defense, by Oskar Morgenstern; The Failure of Atomic Strategy, by F. O. Miksche. Bulletin of the Atomic Scientists, 16(4):139–140, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherwin:1961:BBS

[She61] C. W. Sherwin. Books: The Strategy of Conflict, by Thomas C. Schelling; Nuclear Policy for War and Peace, by Thomas E.

Murray; On War, by Raymond Aron; The Military and Industrial Revolution of Our Time, by Fritz Sternberg. Bulletin of the Atomic Scientists, 17(8):341–342, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sheinin:1964:FSS

[She64a] Yuri Sheinin. Forum: A Soviet scientist looks at disarmament. Bulletin of the Atomic Scientists, 20(1):19–22, January 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sher:1964:LER

[She64b] Lawrence Sher. Letter to the Editor: Reductio ad absurdum. Bulletin of the Atomic Scientists, 20(5):29, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherwin:1964:MSP

[She64c] Chalmers W. Sherwin. The management of science in the public interest. Bulletin of the Atomic Scientists, 20(6):9–12, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [For64].

Sher:1965:LET

[She65] Lawrence Sher. Letter to the Editor: Teach-in twist. Bulletin of the Atomic Scientists, 21(9):23, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sheldon:1969:ASR

[She69] Charles S. Sheldon II. An American “Sputnik” for the Russians? Bulletin of the Atomic Scientists, 25(7):23–25, 27, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shelton:1971:SS

[She71] William R. Shelton. Science in Siberia. Bulletin of the Atomic Scientists, 27(2):23–28, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sheffield:1974:LTE

[She74a] Edwin T. Sheffield. Letter: Throw-away economy. Bulletin of the Atomic Scientists, 30(10):2, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ols74a].

Sheldon:1974:CWT

[She74b] John Sheldon. Communications: To wit, technocracy. Bulletin of the Atomic Scientists, 30(1):3, 48, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha73b].

Shephard:1977:LST

[She77] Frank Shephard. Letter: Science as a tool. Bulletin of the Atomic Scientists, 33(3):3, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bru76].

Sherry:1979:BRB

[She79] Christine Sherry. Book reviews: Controlling Future Arms Trade, by Anne Hessing Cahn, Joseph J. Kruzel, Peter M. Dawkins, and Jacques Huntzinger. Bulletin of the Atomic Scientists, 35(1):43– 44, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sheldon:1980:CWN

[She80] John Sheldon. Commentary: What’s in a name? Bulletin of the Atomic Scientists, 36(8):62, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherry:1982:BRB

[She82] Michael S. Sherry. Book review: National Defense, by James Fallows. Bulletin of the Atomic Scientists, 38(9):39–40, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherry:1983:BRBb

[She83a] Michael S. Sherry. Book review: A History of Strategic Bombing: From the First Hot-Air Balloons to Hiroshima and Nagasaki, by Lee Kennett. Bulletin of the Atomic Scientists, 39(5):42–43, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherry:1983:BRBa

[She83b] Michael S. Sherry. Book review: Strategies of Containment: A Critical Appraisal of Postwar American National Security Policy, by John Lewis Gaddis. Bulletin of the Atomic Scientists, 39(3): 39–40, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shenfield:1985:STA

[She85a] Stephen Shenfield. Soviet thinking about the unthinkable. Bulletin of the Atomic Scientists, 41(2):23–25, February 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shenfield:1985:SMI

[She85b] Stephen Shenfield. Soviets may not imitate Star Wars. Bulletin of the Atomic Scientists, 41(6):38–39, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BS85].

Shermer:1985:LPN

[She85c] Matt Shermer. Letter: Plebiscite on nuclear arms. Bulletin of the Atomic Scientists, 41(2):51–52, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherr:1985:LAC

[She85d] Alan B. Sherr. The languages of arms control. Bulletin of the Atomic Scientists, 41(10):23–29, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherwin:1985:OIN

[She85e] Martin J. Sherwin. Old issues in new editions. Bulletin of the Atomic Scientists, 41(11):40–44, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Swa86].

Sherwin:1985:RHW

[She85f] Martin J. Sherwin. Retrospectives: How well they meant. Bulletin of the Atomic Scientists, 41(7):9–15, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherr:1986:RSW

[She86a] Alan B. Sherr. Removing the Star Wars obstacle. Bulletin of the Atomic Scientists, 42(10):11–13, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherwin:1986:NBS

[She86b] Martin J. Sherwin. Niels Bohr: spurned prophet of arms control. Bulletin of the Atomic Scientists, 42(9):41–45, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherry:1987:PWB

[She87] Michael S. Sherry. Perspective: Was 1945 a break in history? Bulletin of the Atomic Scientists, 43(6):12–15, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sherry:1989:BRB

[She89] Michael S. Sherry. Book review: Of Arms and Men: A History of War, Weapons and Aggression, by Robert L. O’Connell. Bulletin of the Atomic Scientists, 45(7):40–41, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shen:1994:TNW

[She94] Dingli Shen. Toward a nuclear-weapon-free world: a Chinese perspective. Bulletin of the Atomic Scientists, 50(2):51–54, March/ April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shepard:1997:GOS

[She97] William S. Shepard. Guest opinion: Sometimes it is rocket science. Bulletin of the Atomic Scientists, 53(2):4, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shen:2000:WMD

[She00] Dingli Shen. What missile defense says to China. Bulletin of the Atomic Scientists, 56(4):20–21, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/4/20.

Sherman:2005:ACD

[She05a] Jason Sherman. Article commentary: The drone wars. Bulletin of the Atomic Scientists, 61(5):28–37, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/5/28.

Shermer:2005:PPF

[She05b] Michael Shermer. The Pentagon’s psychic friends network. Bulletin of the Atomic Scientists, 61(3):60–61, May/June 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/3/60.

Shen:2007:UDB

[She07] Dingli Shen. Upsetting a delicate balance. Bulletin of the Atomic Scientists, 63(4):37, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/4/37.

Sher:2014:MRC

[She14] Gerson S. Sher. Are medical radioisotopes contributing to global nuclear insecurity? Bulletin of the Atomic Scientists, 70(3):56– 64, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shen:2016:NKN

[She16] Dingli Shen. North Korea, nuclear weapons, and the search for a new path forward. Bulletin of the Atomic Scientists, 72(5):345– ??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1946:AEH

[Shi46a] Edward A. Shils. Atomic energy in the House of Commons. Bulletin of the Atomic Scientists, 1(10):13, 15, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1946:SPI

[Shi46b] Edward A. Shils. Some political implications of the State Dep’t report. Bulletin of the Atomic Scientists, 1(9):7–9, 19, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1947:APS

[Shi47a] Edward A. Shils. American policy and the Soviet ruling group. Bulletin of the Atomic Scientists, 3(9):237–241, 246, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1947:ABV

[Shi47b] Edward A. Shils. The atomic bomb and the veto on sanctions. Bulletin of the Atomic Scientists, 3(2):62–63, February 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). A comment on Prof. P. M. S. Blackett’s criticism of Mr. Baruch’s proposal for the abolition of the veto.

Shils:1947:BAE

[Shi47c] Edward A. Shils. British Atomic Energy Act debate. Bulletin of the Atomic Scientists, 3(2):52–54, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). A report from London of the Parliamentary debate on the British Atomic Energy Act of 1946, and subsequent press comment.

Shils:1947:CPS

[Shi47d] Edward A. Shils. A critique of planning — the Society for Freedom in Science: British scientists organize to stop the trend toward government planning in science. Bulletin of the Atomic Scientists, 3(3):80–82, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1948:FUI

[Shi48a] Edward A. Shils. The failure of the UN AEC: An interpretation. Bulletin of the Atomic Scientists, 4(7):205–210, July 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1948:HLD

[Shi48b] Edward A. Shils. The House of Lords debates international control. Bulletin of the Atomic Scientists, 4(4):122–124, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1948:NPI

[Shi48c] Edward A. Shils. The next phase of the international control discussions. Bulletin of the Atomic Scientists, 4(12):359–362, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1949:BAS

[Shi49] Edward A. Shils. Blackett’s apologia for the Soviet position. Bulletin of the Atomic Scientists, 5(2):34–37, February 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1950:BGC

[Shi50] Edward A. Shils. Books: Grenville Clark’s “A Plan for Peace”. Bulletin of the Atomic Scientists, 6(12):371–, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1951:BEA

[Shi51] Edward Shils. The Bolshevik elite: An analysis of an legend. Bulletin of the Atomic Scientists, 7(3):77–80, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1952:EAP

[Shi52] Edward A. Shils. Editorial: America’s paper curtain. Bulletin of the Atomic Scientists, 8(7):210–217, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Shils:1954:SCT

[Shi54a] Edward Shils. The scientific community: Thoughts after Hamburg. Bulletin of the Atomic Scientists, 10(5):151–155, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ano55p].

Shils:1954:SS

[Shi54b] Edward Shils. A slippery slope. Bulletin of the Atomic Scientists, 10(6):242, 256, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1954:CHR

[Shi54c] Edward A. Shils. Conspiratorial hallucinations: A review of Medford Evans’ The Secret War for the A-Bomb. Bulletin of the Atomic Scientists, 10(2):51–54, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1954:SAP

[Shi54d] Edward A. Shils. Scientists, administrators, and politicians: The report of the Riehlman Committee. Bulletin of the Atomic Scientists, 10(10):371–374, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1955:SSS

[Shi55] Edward Shils. Security and science sacrificed to loyalty. Bulletin of the Atomic Scientists, 11(4):106–109, 130, April 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty.

Shils:1956:MC

[Shi56a] Edward Shils. Milan conference. Bulletin of the Atomic Scientists, 12(2):38–40, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1956:TPP

[Shi56b] Edward A. Shils. Two patterns of publicity, privacy, and secrecy. Bulletin of the Atomic Scientists, 12(6):215–220, June 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1957:FIO

[Shi57] Edward Shils. Freedom and influence: Observations on the scientists’ movement in the United States. Bulletin of the Atomic Scientists, 13(1):13–18, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1961:SDN

[Shi61] Edward Shils. Scientific development in the new states. Bulletin of the Atomic Scientists, 17(2):48–52, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shively:1966:LES

[Shi66] Daniel Shively. Letter to the Editor: On survival. Bulletin of the Atomic Scientists, 22(7):34, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1967:IF

[Shi67] Edward Shils. The intellectuals and the future. Bulletin of the Atomic Scientists, 23(8):7–14, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shinnar:1976:LPS

[Shi76] Reuel Shinnar. Letter: ‘pseudo-scientific’. Bulletin of the Atomic Scientists, 32(2):4, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pre75].

Shinn:1984:RDA

[Shi84] Roger L. Shinn. Review: The days after ‘The Day After’. Bulletin of the Atomic Scientists, 40(2):43–44, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shils:1987:APC

[Shi87] Edward Shils. “America’s paper curtain”. Bulletin of the Atomic Scientists, 43(4):25, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shire:1988:LLE

[Shi88] George B. Shire. Letter: Lasers and enrichment. Bulletin of the Atomic Scientists, 44(4):53, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [CAN88].

Shimabuku:2024:WW

[Shi24] Morgan Shimabuku. Water and war. Bulletin of the Atomic Scientists, 80(4):246–250, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2364499.

Shkarupin:1992:BNP

[Shk92] Vyacheslav Shkarupin. Bulletin: NIMBY PIMBY over nuclear waste. Bulletin of the Atomic Scientists, 48(10):4–5, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shotwell:1946:ABI

[Sho46] James T. Shotwell. The atomic bomb and international organization. Bulletin of the Atomic Scientists, 1(7):8–9, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shore:1976:LC

[Sho76] Lloyd G. Shore. Letter: Not the case. Bulletin of the Atomic Scientists, 32(9):5, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Shobe:1977:LMC

[Sho77] William M. Shobe. Letter: Misplaced confidence. Bulletin of the Atomic Scientists, 33(2):6, 64, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shore:1980:CCN

[Sho80] Lloyd G. Shore. Commentary: Can nuclear weapons be abolished? Bulletin of the Atomic Scientists, 36(10):6–7, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shorr:1989:BWL

[Sho89] David Shorr. Bulletin: Warheads on leave. Bulletin of the Atomic Scientists, 45(8):5, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shore:1992:LON

[Sho92a] Lloyd G. Shore. Letter: Outlawing nuclear weapons. Bulletin of the Atomic Scientists, 48(2):47–48, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel91].

Shorr:1992:BRV

[Sho92b] David Shorr. Book review: The vision thing: The Future Belongs to Freedom, by Eduard Shevardnadze. Bulletin of the Atomic Scientists, 48(1):42–43, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shorr:1992:PNB

[Sho92c] David Shorr. Perspective: NATO: Briefers babble, Bosnia burns. Bulletin of the Atomic Scientists, 48(7):13, 47, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shriver:1970:MHH

[Shr70] Donald W. Shriver, Jr. A memory and a hope: Hiroshima after a quarter century. Bulletin of the Atomic Scientists, 26(7):32– 34, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sag71].

Shrishak:2023:HDA

[Shr23] Kris Shrishak. How to deal with an AI near-miss: Look to the skies. Bulletin of the Atomic Scientists, 79(3):166–169, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2199580.

Stephenson:1987:LSP

[SHT^{+}87] Mike Stephenson, Glenn W. Hawkes, Peter Tewksbury, Claire Thomas, Philip W. Bennett, and Don Manley. Letters: Sex, power, and nuclear language. Bulletin of the Atomic Scientists, 43(7):58–61, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Coh87] and comments [Sta87, Sch87].

Shurcliff:1949:BBW

[Shu49] W. A. Shurcliff. Books: Must We Hide, by R. E. Lapp. Bulletin of the Atomic Scientists, 5(12):352, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shurcliff:1972:CTSb

[Shu72a] William A. Shurcliff. Communications: Technology, science, and the SST. Bulletin of the Atomic Scientists, 28(9):2, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tri72].

Shurcliff:1972:CTSa

[Shu72b] William A. Shurcliff. Communications: Technology, society and SST. Bulletin of the Atomic Scientists, 28(5):2–3, May 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tri71, Tri72].

Shurcliff:1976:ATS

[Shu76] William A. Shurcliff. Active-type solar heating systems for houses: a technology in ferment. Bulletin of the Atomic Scientists, 32(2):30–40, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shull:1978:BRS

[Shu78] Harrison Shull. Books: Research support: the hour glass of American science: The State of Academic Science: The Universities in the Nation’s Research Effort. Volume 1: Summary of Major Findings, by Bruce L. R. Smith and Joseph J. Karlesky. Bulletin of the Atomic Scientists, 34(1):53–54, 56, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shuman:1981:RPE

[Shu81] Michael Shuman. 1980 Rabinowitch Prize Essay on how to eliminate the threat of nuclear war: The mouse that roared. Bulletin of the Atomic Scientists, 37(1):15–22, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Aus81, Pai81a, Swe81].

Shulman:1987:FDI

[Shu87] Marshall D. Shulman. Four decades of irrationality: U.S.–Soviet relations. Bulletin of the Atomic Scientists, 43(9):15–25, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bur88] and erratum [Ano87x].

Sibert:1973:CSS

[Sib73] John Sibert. Communications: A sane solution. Bulletin of the Atomic Scientists, 29(4):54, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, ZSA73a, ZSA73b].

Sich:1996:DS

[Sic96a] Alexander R. Sich. The denial syndrome. Bulletin of the Atomic Scientists, 52(3):38–39, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Hes97, Sta97].

Sich:1996:SQ

[Sic96b] Alexander R. Sich. The shutdown question. Bulletin of the Atomic Scientists, 52(3):36–37, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sich:1996:TWE

[Sic96c] Alexander R. Sich. Truth was an early casualty. Bulletin of the Atomic Scientists, 52(3):32–42, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [HS96].

Sidel:1969:BBS

[Sid69] Victor W. Sidel. Books: The Silent Weapons, by Robin Clarke; Chemical and Biological Warfare, by Seymour M. Hersh. Bulletin of the Atomic Scientists, 25(1):33–34, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Siddiqa:2000:SAN

[Sid00] Ayesha-Agha Siddiqa. South Asia: Nuclear navies? Bulletin of the Atomic Scientists, 56(5):12–14, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/5/12.

Sidhu:2016:SPT

[Sid16] Waheguru Pal Singh (Wps) Sidhu. Seeking a path toward missile nonproliferation. Bulletin of the Atomic Scientists, 72(6):360– 361, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sieger:1952:LEA

[Sie52] John S. Sieger. Letter to the Editor: Attitude of the Catholic Church toward science. Bulletin of the Atomic Scientists, 8(6): 209, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Siegel:1972:YCS

[Sie72] Armand Siegel. The youth culture and the socially concerned scientist. Bulletin of the Atomic Scientists, 28(9):16–19, 21, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sieghart:1980:GNM

[Sie80] Paul Sieghart. Guarding nuclear materials and civil liberties. Bulletin of the Atomic Scientists, 36(5):32–35, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sieghart:1982:NOE

[Sie82] Paul Sieghart. A new objective for energy policy. Bulletin of the Atomic Scientists, 38(1):30–34, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Siegel:1985:BRB

[Sie85] Lenny Siegel. Book review: The Militarization of High Technology, edited by John Tirman. Bulletin of the Atomic Scientists, 41(3):59–60, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tir85].

Siegel:2004:HSG

[Sie04a] Jonas Siegel. The homeland security gravy train. Bulletin of the Atomic Scientists, 60(3):36–37, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/36.

Siegel:2004:NBM

[Sie04b] Jonas Siegel. The new bunker mentality. Bulletin of the Atomic Scientists, 60(6):42–43, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/6/42.

Siegel:2004:SN

[Sie04c] Jonas Siegel. Salvaging nonproliferation. Bulletin of the Atomic Scientists, 60(3):38–39, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/60/3/38.

Siegel:2004:SM

[Sie04d] Jonas Siegel. Still a mystery. Bulletin of the Atomic Scientists, 60(5):68–69, September/October 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/5/68.

Siegel:2005:E

[Sie05a] Jonas Siegel. Be the extremist. Bulletin of the Atomic Scientists, 61(5):71, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/5/71.

Siegel:2005:D

[Sie05b] Jonas Siegel. The determinator. Bulletin of the Atomic Scientists, 61(4):71, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/4/71.

Siegel:2005:INC

[Sie05c] Jonas Siegel. Idaho: Nuclear comes home to roost. Bulletin of the Atomic Scientists, 61(2):18–20, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/2/18.

Siegel:2005:HW

[Sie05d] Jonas Siegel. In Harm’s way. Bulletin of the Atomic Scientists, 61(3):34–35, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/3/34.

Siegel:2005:Z

[Sie05e] Jonas Siegel. In the zone. Bulletin of the Atomic Scientists, 61 (4):38–39, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/4/38.

Siegel:2005:ONN

[Sie05f] Jonas Siegel. The original nuclear nightmare. Bulletin of the Atomic Scientists, 61(5):38–39, September/October 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/5/38.

Siegel:2005:PC

[Sie05g] Jonas Siegel. The paper chase. Bulletin of the Atomic Scientists, 61(5):10, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Siegel:2005:TF

[Sie05h] Jonas Siegel. Time flies. Bulletin of the Atomic Scientists, 61(6): 40–41, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/6/40.

Siegel:2006:BS

[Sie06a] Jonas Siegel. Bugs in the system. Bulletin of the Atomic Scientists, 62(2):34–35, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/2/34.

Siegel:2006:FA

[Sie06b] Jonas Siegel. Farewell to arms. Bulletin of the Atomic Scientists, 62(4):24–25, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/4/24.

Siegel:2006:MG

[Sie06c] Jonas Siegel. Mining the gap. Bulletin of the Atomic Scientists, 62(6):42–43, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/6/42.

Siegel:2006:OG

[Sie06d] Jonas Siegel. The other gas. Bulletin of the Atomic Scientists, 62 (5):24–25, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/5/24.

Siegel:2006:SF

[Sie06e] Jonas Siegel. The sky is filling. Bulletin of the Atomic Scientists, 62(3):32–33, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/3/32.

Siegel:2006:WS

[Sie06f] Jonas Siegel. The way of the spooks. Bulletin of the Atomic Scientists, 62(2):61, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/2/61.

Siegel:2007:PS

[Sie07a] Jonas Siegel. In plane sight. Bulletin of the Atomic Scientists, 63 (1):11, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Siegel:2007:IDE

[Sie07b] Jonas Siegel. Interview: Drew Endy. Bulletin of the Atomic Scientists, 63(3):28–33, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/3/28.

Siegert:2024:HWK

[Sie24] Martin Siegert. How we know Antarctica is rapidly losing more ice. Bulletin of the Atomic Scientists, 80(4):239–245, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2364507.

Simmons:1966:LEW

[SIF66] Henry T. Simmons, David R. Inglis, and Lewis A. Frank. Letter to the Editor: Weapons development in China. Bulletin of the Atomic Scientists, 22(4):29–30, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1985:AAK

[Sig85a] Leon V. Sigal. Antisatellite accord key to summit. Bulletin of the Atomic Scientists, 41(9):16–18, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1985:BRB

[Sig85b] Leon V. Sigal. Book reviews: The Public and Atlantic Defense, edited by Gregory Flynn and Hans Rattinger; European Security and France, by François de Rose, translated by Richard Nice. Bulletin of the Atomic Scientists, 41(9):48–50, October 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1987:CFE

[Sig87a] Leon V. Sigal. Conventional forces in Europe: Signs of a Soviet shift. Bulletin of the Atomic Scientists, 43(10):16–20, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1987:IDF

[Sig87b] Leon V. Sigal. INF deal faces conservation opposition. Bulletin of the Atomic Scientists, 43(4):14–16, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1987:PGS

[Sig87c] Leon V. Sigal. Perspective: Getting over the summit. Bulletin of the Atomic Scientists, 43(1):12–13, January/February 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1988:RCA

[Sig88a] Leon V. Sigal. The Reagan compromise on ABM. Bulletin of the Atomic Scientists, 44(3):10–14, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1988:SNF

[Sig88b] Leon V. Sigal. START nears the finish line. Bulletin of the Atomic Scientists, 44(3):14–15, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1989:NRR

[Sig89] Leon V. Sigal. NATO: Reagan’s radical challenge. Bulletin of the Atomic Scientists, 45(1):38–41, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sigal:1998:JCM

[Sig98] Leon V. Sigal. Jimmy Carter makes a deal. Bulletin of the Atomic Scientists, 54(1):40–46, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gil98].

Sigal:2001:NKH

[Sig01] Leon V. Sigal. North Korea: On hold… again. Bulletin of the Atomic Scientists, 57(3):32–39, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/32.

Sigal:2003:NN

[Sig03] Leon V. Sigal. Negotiating with the North. Bulletin of the Atomic Scientists, 59(6):19–25, November/December 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/6/19.

Sigal:2010:PNK

[Sig10] Leon V. Sigal. Primer — North Korea, South Korea, and the United States: Reading between the lines of the Cheonan attack. Bulletin of the Atomic Scientists, 66(5):35–44, September/ October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
5/35.

Silard:1964:EPA

[Sil64a] John Silard. The economics and politics of arms reduction. Bulletin of the Atomic Scientists, 20(4):6–7, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Silard:1964:FCA

[Sil64b] John Silard. Forum: The case against. Bulletin of the Atomic Scientists, 20(7):18–20, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Silverman:1965:LEA

[Sil65] Jerry Silverman. Letter to the Editor: Art, science. Bulletin of the Atomic Scientists, 21(5):35, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Silard:1966:NWL

[Sil66] John Silard. Nuclear weapons: A liability. Bulletin of the Atomic Scientists, 22(7):15–20, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Silverman:1971:BBC

[Sil71] Albert Silverman. Books: Man and the Computer, by John Diebold; The Bomb and the Computer by Andrew Wilson. Bulletin of the Atomic Scientists, 27(1):50–51, January 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Silver:1979:BRB

[Sil79a] George Silver. Book review: Medicine and the Reign of Technology, by Stanley Joel Reiser. Bulletin of the Atomic Scientists, 35 (4):33–34, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Silver:1979:CLM

[Sil79b] George A. Silver. Commentary: The long march to health. Bulletin of the Atomic Scientists, 35(7):52–54, September 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1946:SRL

[Sim46a] J. A. Simpson. Science and research in the liberated countries. Bulletin of the Atomic Scientists, 1(10):15, May 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1946:SVE

[Sim46b] J. A. Simpson, Jr. A scientist’s visit to England and France. Bulletin of the Atomic Scientists, 1(8):16–17, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1947:SPE

[Sim47] John A. Simpson. The scientists as public educators: A two year summary. Bulletin of the Atomic Scientists, 3(9):243–246, September 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simon:1953:NPB

[Sim53] Franz E. Simon. Nuclear power: a British view. Bulletin of the Atomic Scientists, 9(4):125–128, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1954:BBA

[Sim54a] Alan Simpson. Books: Realities of American Foreign Policy, by George F. Kennan. Bulletin of the Atomic Scientists, 10(9):364– 365, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1954:CCR

[Sim54b] Alan Simpson. Current comment: The re-trial of the Oppenheimer case. Bulletin of the Atomic Scientists, 10(10):387–388, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1954:NN

[Sim54c] Mary M. Simpson. News and notes. Bulletin of the Atomic Scientists, 10(2):59–63, 65, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1954:NNA

[Sim54d] Mary M. Simpson. News and notes: Atomic weapons and U.S. policy; atomic weapons abroad; the Eisenhower plan; atomic disarmament; U.S. atomic power; USAEC; atomic energy abroad; exchange of information with allies; American science; world science; security. Bulletin of the Atomic Scientists, 10(3):106–112, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1954:NRA

[Sim54e] Mary M. Simpson. News roundup: Atomic weapons at home and abroad; the peaceful atom; liberalizing the Atomic Energy Act; nuclear developments abroad; U.S. nuclear developments; security; in the belfry. Bulletin of the Atomic Scientists, 10(4): 141–144, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1955:CCA

[Sim55] Alan Simpson. Current comment: Alternatives to the H-bomb. Bulletin of the Atomic Scientists, 11(1):35–37, January 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1956:LHL

[Sim56] Mary M. Simpson. A long hard look at civil defense: A review of the Holifield Committee hearings. Bulletin of the Atomic Scientists, 12(9):343–348, November 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Fel57].

Simpson:1957:IGY

[Sim57a] John A. Simpson. The International Geophysical Year — a study of our planet. Bulletin of the Atomic Scientists, 13(10):351–356, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1957:RM

[Sim57b] Mary S. Simpson. The race for missiles. Bulletin of the Atomic Scientists, 13(8):302–308, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1959:BLD

[Sim59a] Mary M. Simpson. British Lords debate nuclear disarmament. Bulletin of the Atomic Scientists, 15(6):260–265, June 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1959:CCD

[Sim59b] Mary M. Simpson. Current comment: A divided culture. Bulletin of the Atomic Scientists, 15(7):313–314, September 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kli59].

Simpson:1959:CCF

[Sim59c] Mary M. Simpson. Current comment: The forestalling blow. Bulletin of the Atomic Scientists, 15(9):396–397, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1959:CCTb

[Sim59d] Mary M. Simpson. Current comment: To test or not to test. Bulletin of the Atomic Scientists, 15(10):422–424, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1959:CCTa

[Sim59e] Mary M. Simpson. Current comment: Too many people in the room? Bulletin of the Atomic Scientists, 15(6):272, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1959:SHQ

[Sim59f] Mary M. Simpson. Senator Humphrey’s questions. Bulletin of the Atomic Scientists, 15(8):347–348, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1960:CCSb

[Sim60a] Mary M. Simpson. Current comment: Are space programs worth the cost? Bulletin of the Atomic Scientists, 16(9):373–374, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1960:CCSa

[Sim60b] Mary M. Simpson. Current comment: The scientist in politics: On top or on tap? Bulletin of the Atomic Scientists, 16(1):28– 29, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1960:DAC

[Sim60c] Mary M. Simpson. Defense against CW and BW. Bulletin of the Atomic Scientists, 16(6):254–256, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1961:SAU

[Sim61] J. A. Simpson. The space around us. Bulletin of the Atomic Scientists, 17(5–6):172–174, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1962:BBC

[Sim62] Mary M. Simpson. Books: Science in the Cause of Man, by Gerard Piel. Bulletin of the Atomic Scientists, 18(8):30–31, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1963:SA

[Sim63] Mary S. Simpson. The Snow affair. Bulletin of the Atomic Scientists, 19(4):28–32, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1965:BBS

[Sim65] Mary S. Simpson. Books: Science and the Shabby Curate of Poetry. Essays About the Two Cultures, by Martin Green. Bulletin of the Atomic Scientists, 21(7):31–32, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gil65].

Simpson:1981:SPN

[Sim81] John Alexander Simpson. Some personal notes. Bulletin of the Atomic Scientists, 37(1):26–32, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simmons:1984:WPL

[Sim84a] Adele Simmons. War, peace and liberal education. Bulletin of the Atomic Scientists, 40(10):3s–4s [various between 20 and 21], December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simmons:1984:BRB

[Sim84b] David Simmons. Book review: Workers at Risk: Voices from the Workplace, by Dorothy Nelkin and Michael S. Brown. Bulletin of the Atomic Scientists, 40(7):45, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simon:1984:STC

[Sim84c] Denis F. Simon. Science and technology in China: International influences. Bulletin of the Atomic Scientists, 40(8):5s–7s [various between 24–25], October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simon:1984:BGF

[Sim84d] Julian L. Simon. Bright global future. Bulletin of the Atomic Scientists, 40(9):14–17, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Bar85a, Flo85].

Simon:1984:CSE

[Sim84e] Victor Simon. Commentary: Should Europe declare peace? Bulletin of the Atomic Scientists, 40(6):54, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simon:1985:BRB

[Sim85] Denis Fred Simon. Book review: Chinese Defence Policy, edited by Gerald Segal and William T. Tow. Bulletin of the Atomic Scientists, 41(1):51–53, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simmons:1987:BRB

[Sim87a] Michael Simmons. Book review: Black and Red: W. E. B. Du Bois and the Afro–American Response to the Cold War, 1944– 1963, by Gerald Horne. Bulletin of the Atomic Scientists, 43 (7):56–57, September 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Len88].

Simpson:1987:CDC

[Sim87b] Jeffery Simpson. Canadian defense: Canada roused by military plan. Bulletin of the Atomic Scientists, 43(8):9–10, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gol88d].

Simon:1988:BRB

[Sim88] Denis Fred Simon. Book review: The Chinese Army after Mao, by Ellis Joffe. Bulletin of the Atomic Scientists, 44(5):46–48, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1990:NDA

[Sim90a] John Simpson. Nonproliferation’s divided agenda. Bulletin of the Atomic Scientists, 46(6):17, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1990:NRN

[Sim90b] John Simpson. NPT review: Nonproliferation agenda beyond 1990. Bulletin of the Atomic Scientists, 46(6):38–39, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1991:PNS

[Sim91] John Simpson. Perspective: NPT stronger after Iraq. Bulletin of the Atomic Scientists, 47(8):12–13, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simpson:1995:GOC

[Sim95] John A. Simpson. Guest opinion: A challenge for the 21st century. Bulletin of the Atomic Scientists, 51(6):4, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sims:2011:STC

[Sim11] Nicholas A. Sims. A simple treaty, a complex fulfillment: a short history of the biological weapons convention review Conferences. Bulletin of the Atomic Scientists, 67(3):8–15, May/June 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/3/8.

Simo:2016:WIT

[Sim16] Augustin Simo. What if there’s a next time? Preparedness after Chernobyl and Fukushima. Bulletin of the Atomic Scientists, 72 (4):265–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sinnott:1956:PVS

[Sin56] Edmund W. Sinnott. Part VII: Science and religion: Science and the human spirit. Bulletin of the Atomic Scientists, 12(10):360– 364, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1961:DS

[Sin61] J. David Singer. Deterrence and shelters. Bulletin of the Atomic Scientists, 17(8):310–315, October 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1963:FMD

[Sin63a] J. David Singer. Forum: Military doctrine. Bulletin of the Atomic Scientists, 19(8):35, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1963:LED

[Sin63b] J. David Singer. Letter to the Editor: Disenchantment with diplomacy. Bulletin of the Atomic Scientists, 19(2):35, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1963:PRP

[Sin63c] J. David Singer. Peace research, peace action. Bulletin of the Atomic Scientists, 19(1):13–17, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and response [CS63].

Singh:1965:SRI

[Sin65] Jagjit Singh. Scientific research in India. Bulletin of the Atomic Scientists, 21(2):41–43, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1967:BBS

[Sin67a] Maxine Singer. Books: The MIT Symposium on American Women in Science and Engineering, edited by J. A. Mattfeld and C. G. Van Aken. Bulletin of the Atomic Scientists, 23(5): 57–58, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sinsheimer:1967:EB

[Sin67b] Robert Sinsheimer. The end of the beginning. Bulletin of the Atomic Scientists, 23(2):8–12, February 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1970:ESS

[Sin70] S. Fred Singer. Exploring space in the seventies. Bulletin of the Atomic Scientists, 26(9):22–23, November 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1976:BFF

[Sin76] S. Fred Singer. Books: The facts and fancies of the ‘conservation ethic’ Mineral Resources and the Environment. Bulletin of the Atomic Scientists, 32(6):39–41, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1978:BBS

[Sin78] S. Fred Singer. Books: The State of Competition in the Coal Industry. Bulletin of the Atomic Scientists, 34(7):51, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1979:SCC

[Sin79a] Grace L. Singer. Science and the citizen: a case study in Hudson County: Citizens defend the urban coast. Bulletin of the Atomic Scientists, 35(6):47–52, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1979:CCU

[Sin79b] J. David Singer. Commentary: Campaign for a U.S. peace academy. Bulletin of the Atomic Scientists, 35(3):84–86, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1981:BRB

[Sin81] Grace L. Singer. Book review: The Reserve Mining Controversy: Science, Technology and Environmental Quality, by Robert V. Bartlett. Bulletin of the Atomic Scientists, 37(6):40–42, June/ July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sinden:1983:CIL

[Sin83] Frank W. Sinden. Commentary: Inventing the low energy house. Bulletin of the Atomic Scientists, 39(9):56–58, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sinclair:1991:LCM

[Sin91] John Sinclair. Letter: Cover misleading. Bulletin of the Atomic Scientists, 47(6):45, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Singer:1992:WTN

[Sin92] S. Fred Singer. Warming theories need warning label. Bulletin of the Atomic Scientists, 48(5):34–39, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MJD92].

Singer:2005:ONS

[Sin05a] Ron Singer. Oil: Nigeria’s slippery politics. Bulletin of the Atomic Scientists, 61(1):12–16, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/1/12.

Singh:2005:ACB

[Sin05b] Jai Singh. Arms control: Bomb-grade bargain. Bulletin of the Atomic Scientists, 61(6):16–18, November/December 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/6/16.

Sinovets:2014:WWR

[Sin14] Polina Sinovets. Women and weapons: Redressing the gender gap:o a Ukrainian response. Bulletin of the Atomic Scientists, 70 (5):21–23, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sinovets:2017:BBB

[Sin17] Polina Sinovets. Ban the bomb by … banning the bomb? Bulletin of the Atomic Scientists, 73(3):197–198, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1315090.

Sipos:1986:LTB

[Sip86] István Sipos. Letter: UN test ban resolutions. Bulletin of the Atomic Scientists, 42(5):48, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sivard:1975:LTE

[Siv75] Rith Leger Sivard. Let them eat bullets! A statistical portrait of world militarism. Bulletin of the Atomic Scientists, 31(4):6–10, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sivard:1976:BFF

[Siv76] Ruth Leger Sivard. Books: A feast of facts: Arms Uncontrolled, by Frank Burnaby and Ronald Huisken. Bulletin of the Atomic Scientists, 32(5):64, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sivard:1980:BRB

[Siv80] Ruth Leger Sivard. Book review: The Price of Defense: A New Strategy for Military Spending, by The Boston Study Group. Bulletin of the Atomic Scientists, 36(4):46, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stagner:1964:LEP

[SJ64] Ross Stagner and Nehemiah Jordan. Letter to the Editor: Is psychology irrelevant? Bulletin of the Atomic Scientists, 20(3): 24–25, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Jor63].

Smith:1968:RSP

[SJ68] Philip M. Smith and Rodney W. Johnson. Reports: From the South Pole to the Moon: Parallels in exploration. Bulletin of the Atomic Scientists, 24(10):35–37, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Seckler:1982:SWM

[SJ82] David Seckler and Deep Joshi. Sukhomajri: water management in India. Bulletin of the Atomic Scientists, 38(3):26–30, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Slomich:1969:SPP

[SK69] Sidney J. Slomich and Robert E. Kantor. Social psychopathology of political assassination. Bulletin of the Atomic Scientists, 25(3): 9–12, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stewart:1984:CME

[SK84] A. W. Stewart and G. W. Kneale. Commentary: Mortality experiences of A-bomb survivors. Bulletin of the Atomic Scientists, 40 (5):61–62, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinberg:1985:LIA

[SK85] Gerald M. Steinberg and Wayland Kennet. Letter: Irrelevant attack on Israel. Bulletin of the Atomic Scientists, 41(11):64– 65, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ken85].

Stanley:2003:NCT

[SK03] Richard Stanley and Michael Ryan Kraig. The NPT: Can this treaty be saved? Bulletin of the Atomic Scientists, 59(5):59–65, September/October 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/59/5/59.

Skelly:1990:CSS

[Ske90] James M. Skelly. “Civil society” and the state. Bulletin of the Atomic Scientists, 46(2):19, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skinner:1948:WHE

[Ski48] H. W. B. Skinner. The work of the Harwell Establishment. Bulletin of the Atomic Scientists, 4(4):107–109, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skinner:1953:AEP

[Ski53] H. W. B. Skinner. Atomic energy and private enterprise. Part II. atomic energy in Postwar Britain. Bulletin of the Atomic Scientists, 9(5):156–161, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skobeltzyn:1958:CRE

[Sko58] D. V. Skobeltzyn. A chain reaction of errors. Bulletin of the Atomic Scientists, 14(7):246–249, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skolnikoff:1967:BDI

[Sko67] Eugene B. Skolnikoff. Birth and death of an idea: Research in aid. Bulletin of the Atomic Scientists, 23(7):38–40, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skolnikoff:1968:SPP

[Sko68] Eugene Skolnikoff. Science and public policy meeting at AAAS. Bulletin of the Atomic Scientists, 24(6):49, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skubitz:1972:CSM

[Sku72] Joe Skubitz. Communications: Salt mine. Bulletin of the Atomic Scientists, 28(3):4, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lew71a].

Stewart:1946:ABP

[SL46] Irving Stewart and Captain Lavendar. Atomic bomb patents. Bulletin of the Atomic Scientists, 1(7):10–11, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schubert:1958:RMB

[SL58] Jack Schubert and Ralph E. Lapp. Radiation and man: Basic data: Global radiation limits. Bulletin of the Atomic Scientists, 14(1):23–26, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Skinner:1980:CCN

[SL80] L. Walter Skinner and John R. Lamarsh. Commentary: On China’s nuclear power program. Bulletin of the Atomic Scientists, 36(9):60, 62, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lam80].

Smith:1983:CWU

[SL83] Wayne S. Smith and Abraham F. Lowenthal. Commentary: Why the United States is in Central America. Bulletin of the Atomic Scientists, 39(10):58–60, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ken83].

Seaborg:1987:GEM

[SL87] Glenn T. Seaborg and Benjamin S. Loeb. Guest editorial: Make the partial test ban comprehensive. Bulletin of the Atomic Scientists, 43(4):3, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salisbury:2013:TCS

[SL13] Daniel Salisbury and David Lowrie. Targeted: a case study in Iranian illicit missile procurement. Bulletin of the Atomic Scientists, 69(3):23–30, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/3/23.

Slawson:1976:LMB

[Sla76] Barbara M. Slawson. Letter: Mind-boggling. Bulletin of the Atomic Scientists, 32(9):7, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Slater:1988:BRB

[Sla88] Jerome Slater. Book review: National Security and United States Policy toward Latin America, by Lars Schoultz. Bulletin of the Atomic Scientists, 44(5):50–52, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1952:CDN

[Sle52] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 8(9):313–314, December 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNa

[Sle53a] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(1):26, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNb

[Sle53b] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(2):60–61, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNc

[Sle53c] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(3):95–96, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNd

[Sle53d] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(5):199, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNe

[Sle53e] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(6):232, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNf

[Sle53f] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(8):315, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1953:CDNg

[Sle53g] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 9(10):388, December 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1954:CDNb

[Sle54a] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 10(2):58, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1954:CDNc

[Sle54b] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 10(3):105, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1954:CDNe

[Sle54c] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 10(5):172, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sleevi:1954:CDNh

[Sle54d] Eugene J. Sleevi. Civil defense news. Bulletin of the Atomic Scientists, 10(8):331, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Slessor:1956:GDL

[Sle56] Sir John Slessor. The great deterrent and its limitations. Bulletin of the Atomic Scientists, 12(5):140–146, May 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Slemrod:1984:CEN

[Sle84] Joel Slemrod. Commentary: The economics of nuclear fear. Bulletin of the Atomic Scientists, 40(4):42–43, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sobel:1987:LPM

[SLM87] Michael I. Sobel, Steven R. Lampman, and Gerald E. Marsh. Letters: Protecting missiles with SDI. Bulletin of the Atomic Scientists, 43(4):54–, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar87a].

Slomich:1968:ACD

[Slo68] Sidney J. Slomich. Arms control and disarmament: The great evasion. Bulletin of the Atomic Scientists, 24(6):18–22, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Slovic:2012:PGR

[Slo12] Paul Slovic. The perception gap: Radiation and risk. Bulletin of the Atomic Scientists, 68(3):67–75, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/68/3/67.

Sly:1999:SP

[Sly99] Liz Sly. A state of paranoia. Bulletin of the Atomic Scientists, 55(5):38–43, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/5/38.

Somerville:1981:CFM

[SM81] John Somerville and Michael Myerson. Commentary: Four minutes to midnight. Bulletin of the Atomic Scientists, 37(4):58, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel81d].

Scheer:1982:LB

[SM82] Robert Scheer and Robert McNamara. Looking back at the 1960s. Bulletin of the Atomic Scientists, 38(6):18, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szegedy-Maszak:1989:IB

[SM89a] Marianne Szegedy-Maszak. Inside the Beltway. Bulletin of the Atomic Scientists, 45(1):20, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mac89].

Szegedy-Maszak:1989:MRF

[SM89b] Marianne Szegedy-Maszak. The movement: Rise and fall of the Washington peace industry. Bulletin of the Atomic Scientists, 45(1):18–23, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Saleska:1990:HCE

[SM90] Scott Saleska and Arjun Makhijani. Hanford cleanup: explosive solution. Bulletin of the Atomic Scientists, 46(8):14–16, 18–20, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Speed:2005:DD

[SM05] Roger Speed and Michael May. Dangerous doctrine. Bulletin of the Atomic Scientists, 61(2):38–49, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/2/38.

Smith:2014:TIC

[SM14] Matt Smith and Katharine Mieszkowski. Treasure Island cleanup exposes Navy’s mishandling of its nuclear past. Bulletin of the Atomic Scientists, 70(3):65–78, May 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smalley:1976:LJG

[Sma76a] Walter B. Smalley. Letter: ‘Jews, God and country’. Bulletin of the Atomic Scientists, 32(2):3, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bec75].

Smalley:1976:LRB

[Sma76b] Walter B. Smalley. Letter: Refined barbarism. Bulletin of the Atomic Scientists, 32(7):3, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gan76].

Smalley:1977:LB

[Sma77] Walter B. Smalley. Letter: Down to basics. Bulletin of the Atomic Scientists, 33(5):2, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wil77a].

Sandberg:2008:HCW

[SMĆ08] Anders Sandberg, Jason Matheny, and Milan Ćirković. How can we reduce the risk of human extinction? Bulletin of the Atomic Scientists, 64(4):??, September 9, 2008. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://newstrust.net/stories/25990;   http://www.
thebulletin.org/node/4317;   http://www.thebulletin.org/
web-edition/features/how-can-we-reduce-the-risk-of-human-

extinction. Special report.

Smedley:1974:CUB

[Sme74] Frederic C. Smedley. Communications: Update the Baruch plan. Bulletin of the Atomic Scientists, 30(1):2, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smedley:1976:LDU

[Sme76] Frederic C. Smedley. Letter: Defiance of the U.N. Bulletin of the Atomic Scientists, 32(3):2, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smeets:2022:UHC

[Sme22] Max Smeets. A US history of not conducting cyber attacks. Bulletin of the Atomic Scientists, 78(4):208–213, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2087380.

Smith:1953:WCC

[Smi53] O. G. Smith. Warning and communications in civil defense. Bulletin of the Atomic Scientists, 9(7):283–284, September 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1954:BBAf

[Smi54a] Cyril Smith. Books: An American In Europe, by Egon Larsen. Bulletin of the Atomic Scientists, 10(6):238, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1954:BBC

[Smi54b] Cyril S. Smith. Books: Snow Crystals — Natural and Artificial, by Ukichiro Nakaya. Bulletin of the Atomic Scientists, 10(10): 394, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1954:BBAa

[Smi54c] Cyril Stanley Smith. Books: Report on the Atom. What You Should Know About the Atomic Energy Program of the United States, by Gordon Dean. Bulletin of the Atomic Scientists, 10(1): 28–30, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1955:RGS

[Smi55] Alice K. Smith. Report from Geneva: Sidelights on Geneva. Bulletin of the Atomic Scientists, 11(8):276–277, 296, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1958:CCa

[Smi58a] Alice K. Smith. Current comment. Bulletin of the Atomic Scientists, 14(3):123–124, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1958:CCb

[Smi58b] Alice K. Smith. Current comment. Bulletin of the Atomic Scientists, 14(4):151–152, April 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1958:CCG

[Smi58c] Alice K. Smith. Current comment: “The gravest problem … ’. Bulletin of the Atomic Scientists, 14(5):193–194, May 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1958:BDU

[Smi58d] Alice Kimball Smith. Behind the decision to use the atomic bomb: Chicago 1944–45. Bulletin of the Atomic Scientists, 14(8):288– 312, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1958:BBF

[Smi58e] Marianne Smith. Books: Faster, Faster, by W. J. Eckert and Rebecca Jones. Bulletin of the Atomic Scientists, 14(5):196, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1958:BBM

[Smi58f] Marianne Smith. Books: The Modern Universe, by Raymond A. Lyttleton; Atoms and the Universe, by G. O Jones, J. Rotblat, and G. J. Whitrow. Bulletin of the Atomic Scientists, 14(5):196, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1959:BBM

[Smi59a] Alice Kimball Smith. Books: Command the Morning, by Pearl S. Buck. Bulletin of the Atomic Scientists, 15(9):395, 400, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1959:BBF

[Smi59b] Cyril Stanley Smith. Books: Accent on Form. An Anticipation of the Science of Tomorrow, by Lancelot Law Whyte. Bulletin of the Atomic Scientists, 15(2):92, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1960:BBB

[Smi60] Alice K. Smith. Books: Communism and British Intellectuals, by Neal Wood. Bulletin of the Atomic Scientists, 16(2):77, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1961:BBM

[Smi61] Cyril S. Smith. Books: Concepts of Mass in Classical and Modern Physics, by Max Jammer. Bulletin of the Atomic Scientists, 17 (10):437, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1963:FFB

[Smi63] Alice K. Smith. Forum: Fight over a bomb. Bulletin of the Atomic Scientists, 19(8):33, October 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1965:PHS

[Smi65] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. University of Chicago Press, Chicago, IL, USA and London, UK, 1965. xiv + 591 pp. LCCN Q127.U6 S6.

Smith:1967:RPG

[Smi67] Jean Edward Smith. Red Prussianism of the German Democratic Republic. Bulletin of the Atomic Scientists, 23(5):24–30, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1968:BBP

[Smi68] Alice Kimball Smith. Books: Manhattan Project: The Untold Story of the Making of the Atomic Bomb, by Stephane Groueff. Bulletin of the Atomic Scientists, 24(1):52–53, January 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1969:CCW

[Smi69a] Lewis V. Smith. Communiques: Canada’s water rights. Bulletin of the Atomic Scientists, 25(4):38, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1969:PIC

[Smi69b] Philip M. Smith. Prospects for international cooperation on the Moon: The Antarctic analogy. Bulletin of the Atomic Scientists, 25(7):36–40, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1970:AFA

[Smi70a] Alice Kimball Smith. Los Alamos: Focus of an age. Bulletin of the Atomic Scientists, 26(6):15–20, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1970:PHS

[Smi70b] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-69026-8. xvi + 398 pp. LCCN Q127.U6 S6 1971.

Smith:1970:ICA

[Smi70c] Philip M. Smith. International cooperation in Antarctica — the next decade. Bulletin of the Atomic Scientists, 26(10):29–32, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Smith:1971:RES

[Smi71] Delbert D. Smith. Reports: Educational satellite telecommunication: The challenge of a new technology. Bulletin of the Atomic Scientists, 27(4):14–18, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1972:HSF

[Smi72a] Arthur Smith. How the space farers fare: A British view of American and Russian programs since Apollo 11. Bulletin of the Atomic Scientists, 28(4):18–23, 26, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1972:CMT

[Smi72b] David Smith. Communications: On “The mounting tide of unreason”. Bulletin of the Atomic Scientists, 28(2):2–3, February 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab71c, Man72].

Smith:1972:LG

[Smi72c] Joseph V. Smith. I. Lunar geology. Bulletin of the Atomic Scientists, 28(4):31–36, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1973:RLR

[Smi73] Waldo E. Smith. Resources and long-range forecasts. Bulletin of the Atomic Scientists, 29(5):21–23, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hol73].

Smith:1974:BRA

[Smi74a] A. Robert Smith. The breeder reactor: Another SST? Bulletin of the Atomic Scientists, 30(8):12–13, October 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1974:NSE

[Smi74b] A. Robert Smith. No shortage of energy lobbying. Bulletin of the Atomic Scientists, 30(5):11–13, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1974:LPS

[Smi74c] David Smith. Letters: Poverty of scientific inquiry. Bulletin of the Atomic Scientists, 30(5):4–5, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pla73, Sag73].

Smith:1974:LSC

[Smi74d] Marianne F. Smith. Letter: A single coin. Bulletin of the Atomic Scientists, 30(9):48, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pas74].

Smith:1975:ENG

[Smi75a] A. Robert Smith. ERDA: The new glamour agency: Congress seeks a broader base for energy research. Bulletin of the Atomic Scientists, 31(1):29–31, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1975:NSA

[Smi75b] A. Robert Smith. The new scientist–advocates. Bulletin of the Atomic Scientists, 31(2):16–17, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Yeo75].

Smith:1975:BBP

[Smi75c] Alice Kimball Smith. Books: Toward a Political Sociology of Science, by Stuart S. Blume. Bulletin of the Atomic Scientists, 31(1):45–46, January 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1975:LWN

[Smi75d] Anthony Wayne Smith. Letters: We need restraint. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Smith:1976:LDH

[Smi76] Philip B. Smith. Letter: Delay in Holland. Bulletin of the Atomic Scientists, 32(3):3, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [SS76].

Smil:1977:IET

[Smi77a] Vaclac Smil. Intermediate energy technology in China. Bulletin of the Atomic Scientists, 33(2):25–31, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1977:WRP

[Smi77b] Cyril Stanley Smith. Weapons are the real problem. Bulletin of the Atomic Scientists, 33(3):69, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smil:1979:REH

[Smi79] Vaclav Smil. Renewable energies: how much and how renewable? Bulletin of the Atomic Scientists, 35(10):12–19, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SS80].

Smith:1980:CLH

[Smi80] David Smith. Commentary: On ‘learning how to live … ’. Bulletin of the Atomic Scientists, 36(3):62, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mat80b].

Smith:1981:CRO

[Smi81] Richard K. Smith. Commentary: Remember Osirak. Bulletin of the Atomic Scientists, 37(7):63–64, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smil:1982:EMC

[Smi82a] Vaclav Smil. Ecological mismanagement in China. Bulletin of the Atomic Scientists, 38(8):18–23, October 1982. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hum83].

Smith:1982:SPI

[Smi82b] Alice Kimball Smith. Scientists and public issues. Bulletin of the Atomic Scientists, 38(10):38–45, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1983:BRB

[Smi83] Gaddis Smith. Book reviews: Life after Doomsday: A Survivalist Guide to Nuclear War and Other Major Decisions, by Bruce D. Clayton; Survive the Coming Nuclear War: How to Do It, by Ronald L. Cruit and Robert L. Cruit; Nuclear War Survival Skills, by Cresson H. Kearney; The Nuclear Survival Handbook, by Barry Popkens. Bulletin of the Atomic Scientists, 39(6):29– 30, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [RS83].

Smith:1984:HRI

[Smi84a] David Smith. On ‘Human rights imperialism’. Bulletin of the Atomic Scientists, 40(2):56, 58, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1984:ACD

[Smi84b] Gerard C. Smith. The Arms Control and Disarmament Agency: an unfinished history. Bulletin of the Atomic Scientists, 40(4): 13–17, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Loe84].

Smith:1984:CDD

[Smi84c] William W. Smith III. Commentary: In defense of disarmament. Bulletin of the Atomic Scientists, 40(8):48–49, October 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra84a].

Smith:1985:LWC

[Smi85a] Blackwell Smith. Letter: We can stop the nuclear arms race. Bulletin of the Atomic Scientists, 41(6):56–57, June/July 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1985:LDV

[Smi85b] David Smith. Letter: Different view of Trudeau. Bulletin of the Atomic Scientists, 41(5):52, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tru85].

Smith:1985:NDE

[Smi85c] Gerard C. Smith. No dead end for arms control. Bulletin of the Atomic Scientists, 41(1):3–4, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1985:BRB

[Smi85d] Joseph V. Smith. Book review: The Cold and the Dark: The World after Nuclear War, by Paul R. Ehrlich, Carl Sagan, Donald Kennedy and Walter Orr Roberts. Bulletin of the Atomic Scientists, 41(1):49–51, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1987:PPM

[Smi87a] Blackwell Smith. Perspective: Priorities in a military culture. Bulletin of the Atomic Scientists, 43(5):9–10, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1987:LST

[Smi87b] Jeff Smith. Letter: Star Trek/Star Wars connection. Bulletin of the Atomic Scientists, 43(6):60, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom87].

Smith:1987:RSW

[Smi87c] Jeff Smith. Reagan, Star Wars, American culture. Bulletin of the Atomic Scientists, 43(1):19–25, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Swe87a, Gom87, Kel87, You87b].

Smith:1988:BRB

[Smi88] Jeff Smith. Book review: Nuclear Fear: A History of Images, by Spencer R. Weart. Bulletin of the Atomic Scientists, 44(9):42– 43, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1989:ACC

[Smi89] Gerard C. Smith. Arms control criteria: The Gorbachev spin. Bulletin of the Atomic Scientists, 45(4):35–37, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smil:1990:BRH

[Smi90a] Vaclav Smil. Book review: Heat, little light Global Warming: Are We Entering the Greenhouse Century?, by Stephen H. Schneider. Bulletin of the Atomic Scientists, 46(5):39, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kel90].

Smith:1990:PTN

[Smi90b] Gerard C. Smith. Perspective: Take nuclear weapons into custody. Bulletin of the Atomic Scientists, 46(10):12–13, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jon91b].

Smith:1990:LAA

[Smi90c] Lucius Smith III. Letter: Andy Atom not in Mickey’s league. Bulletin of the Atomic Scientists, 46(4):54, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lan90d].

Smithson:1991:CIO

[Smi91] Amy E. Smithson. Chemical inspectors: On the outside looking in? Bulletin of the Atomic Scientists, 47(8):22–25, October 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DeV92].

Smithson:1992:CWE

[Smi92a] Amy E. Smithson. Chemical weapons: The end of the beginning. Bulletin of the Atomic Scientists, 48(8):36–40, October 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1992:NMP

[Smi92b] Amy E. Smithson. No more poison bullets. Bulletin of the Atomic Scientists, 48(8):37–38, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1992:OSR

[Smi92c] Amy E. Smithson. Open Skies ready for takeoff. Bulletin of the Atomic Scientists, 48(1):17–21, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1992:RCW

[Smi92d] Amy E. Smithson. Report: Chemical weapons: Tottering toward a treaty. Bulletin of the Atomic Scientists, 48(6):9–11, July/ August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1993:BP

[Smi93a] Amy E. Smithson. Basic provisions. Bulletin of the Atomic Scientists, 49(3):42–43, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1993:CDW

[Smi93b] Amy E. Smithson. Chemicals destruction: The work begins. Bulletin of the Atomic Scientists, 49(3):38–43, April 1993. See comment [Wil93].

Smithson:1993:PC

[Smi93c] Amy E. Smithson. The Preparatory Commission. Bulletin of the Atomic Scientists, 49(3):40, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1993:RCW

[Smi93d] Amy E. Smithson. Report: Chemical weapons: Conventional wait. Bulletin of the Atomic Scientists, 49(7):10–11, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1994:LSD

[Smi94a] Nancy Smith. Letter: Stacked deck. Bulletin of the Atomic Scientists, 50(1):63, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Abb93].

Smithson:1994:PRW

[Smi94b] Amy E. Smithson. Perspective: Russia wants plastics, too. Bulletin of the Atomic Scientists, 50(3):14–15, May/June 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1995:GOH

[Smi95a] Wayne S. Smith. Guest opinion: Help Cuba with nuclear power. Bulletin of the Atomic Scientists, 51(5):4, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Hug95].

Smithson:1995:PIC

[Smi95b] Amy E. Smithson. Perspective: Incineration confrontation. Bulletin of the Atomic Scientists, 51(1):14–15, 67, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [WS95].

Smith:1996:LNP

[Smi96] Nancy Smith. Letter: No plan for abolition. Bulletin of the Atomic Scientists, 52(3):60, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cor96].

Smith:1997:BRP

[Smi97] Philip Smith. Book review: Pugwash, thinly disguised: Menachem’s Seed, by Carl Djerassi. Bulletin of the Atomic Scientists, 53(6):57–58, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithson:1998:SET

[Smi98] Amy E. Smithson. Guest opinion: Senate emasculates treaty. Bulletin of the Atomic Scientists, 54(4):6, July/August 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:2000:SW

[Smi00] Jeff Smith. Star Wars. Bulletin of the Atomic Scientists, 56(5): 61–63, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/56/5/61.

Smith:2005:EMD

[Smi05] Gayle E. Smith. An epidemic of mass destruction. Bulletin of the Atomic Scientists, 61(6):80, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/61/6/80.

Smith:2013:CCC

[Smi13] Timothy M. Smith. Climate change: Corporate sustainability in the supply chain. Bulletin of the Atomic Scientists, 69(3):43–52, May/June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/69/
3/43.

Smithson:2016:WCT

[Smi16] Amy E. Smithson. Why cooperative threat reduction still matters — especially for biological dangers. Bulletin of the Atomic Scientists, 72(5):322–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smithberger:2021:USN

[Smi21] Mandy Smithberger. The United States needs to cut military spending and shift money to two pressing threats: Pandemics and climate change. Bulletin of the Atomic Scientists, 77(5): 251–254, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1964248.

Smoke:1989:LWM

[Smo89] Richard Smoke. Letter: Warring minds. Bulletin of the Atomic Scientists, 45(7):44, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ede89].

Sterman:2022:DWB

[SMRVS22] John Sterman, William Moomaw, Juliette N. Rooney-Varga, and Lori Siegel. Does wood bioenergy help or harm the climate? Bulletin of the Atomic Scientists, 78(3):128–138, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2062933.

Smyth:1950:PSF

[Smy50a] Henry D. Smyth. The place of science in a free society. Bulletin of the Atomic Scientists, 6(6):180–185, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smyth:1950:RNL

[Smy50b] Henry D. Smyth. The role of the National Laboratories in atomic energy development. Bulletin of the Atomic Scientists, 6(1):5–8, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smyth:1951:SRS

[Smy51] Henry D. Smyth. The stockpile and rationing of scientific manpower. Bulletin of the Atomic Scientists, 7(2):38–42, 64, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smyth:1954:DO

[Smy54] Henry DeWolf Smyth. Dissenting opinion. Bulletin of the Atomic Scientists, 10(7):280–282, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smyth:1960:ICJ

[Smy60] Henry Smyth. International control in jeopardy? Bulletin of the Atomic Scientists, 16(5):161, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smythe:1961:CS

[Smy61] Dallas W. Smythe. Communications satellites. Bulletin of the Atomic Scientists, 17(2):65–70, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smythe:1964:LEO

[Smy64] Dallas W. Smythe. Letter to the Editor: Orbiting bombs and the MOL [Manned Orbiting Laboratories]. Bulletin of the Atomic Scientists, 20(8):37–38, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spo64].

Schulman:1989:LRD

[SN89] Ephraim Schulman and Robert S. Norris. Letter: Recount demanded. Bulletin of the Atomic Scientists, 45(7):42, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NA89k].

Stevenson:1954:VDS

[SNB^{+}54] Adlai Stevenson, Richard M. Nixon, Chester Bowles, John Foster Dulles, Arthur Radford, and Dwight D. Eisenhower. Verbal diplomacy: A second look at the “New Look”. Bulletin of the Atomic Scientists, 10(5):148–150, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snell:1957:BBP

[Sne57] Arthur H. Snell. Books: The Prospects of Nuclear Power and Technology, by Gerald Wendt. Bulletin of the Atomic Scientists, 13(10):372, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1968:BBC

[Sno68a] Joel A. Snow. Books: A Comprehensible World, by Jeremy Bernstein. Bulletin of the Atomic Scientists, 24(7):26–27, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1968:BBP

[Sno68b] Joel A. Snow. Books: The Politics of Pure Science, by Daniel S. Greenberg. Bulletin of the Atomic Scientists, 24(5):34–36, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1968:I

[Sno68c] Joel Alan Snow. An introduction. Bulletin of the Atomic Scientists, 24(8):24–25, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1969:BBK

[Sno69] Joel A. Snow. Books: Public Knowledge: The Social Dimension of Science, by John Ziman. Bulletin of the Atomic Scientists, 25 (3):36–37, 48, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1980:MML

[Sno80] Donald M. Snow. MX: Maginot line of the 1980s. Bulletin of the Atomic Scientists, 36(9):21–25, November 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1981:CDN

[Sno81] Donald M. Snow. Commentary: Demise of nuclear deterrence? Bulletin of the Atomic Scientists, 37(9):62–64, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar81a] and erratum [Ano82j].

Snyder:1985:DFI

[Sny85] Edward F. Snyder. Deterrence: from fear to interdependence. Bulletin of the Atomic Scientists, 41(9):40–42, October 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shelley:2006:CA

[SO06] Louise Shelley and Robert Orttung. Criminal acts. Bulletin of the Atomic Scientists, 62(5):22–23, September/October 2006. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/62/5/22.

SoaresPetrarolha:1996:RSA

[Soa96] Fabio Lacerda Soares Petrarolha. Report: South America: Rivals to march side by side. Bulletin of the Atomic Scientists, 52(5): 14, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sobczyk:1991:LDH

[Sob91a] Garret Sobczyk. Letter: Dealing with hijackers. Bulletin of the Atomic Scientists, 47(3):47, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sobell:1991:LIR

[Sob91b] Morton Sobell. Letter: Iraq and the Rosenbergs. Bulletin of the Atomic Scientists, 47(4):47, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sobel:2021:MTG

[Sob21] Adam Sobel. Making the transition to a green economy: What is our responsibility as citizens? Bulletin of the Atomic Scientists, 77(2):67–69, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1885821.

Socolar:1954:BBE

[Soc54] Sidney J. Socolar. Books: Atomic Energy — A Survey, edited by J. Rotblat. Bulletin of the Atomic Scientists, 10(9):364, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Socolar:1955:LEU

[Soc55] Sidney J. Socolar. Letter to the Editor: [U.N. Disarmament Commission Subcommittee meeting]. Bulletin of the Atomic Scientists, 11(7):271, 273, September 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Socolow:2007:FNU

[Soc07] Robert H. Socolow. Facing new unknowns. Bulletin of the Atomic Scientists, 63(1):45–46, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/1/45.

Socolow:2015:CCD

[Soc15] Robert Socolow. Climate change and destiny studies: Creating our near and far futures. Bulletin of the Atomic Scientists, 71 (6):18–28, November 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Socolow:2020:CCC

[Soc20] Robert H. Socolow. Contending with climate change: The next 25 years. Bulletin of the Atomic Scientists, 76(6):294–301, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1846410.

Smetana:2021:DGN

[SOE21] Michal Smetana, Michal Onderco, and Tom Etienne. Do Germany and The Netherlands want to say goodbye to US nuclear weapons? Bulletin of the Atomic Scientists, 77(4):215–221, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1941603.

Sohn:1961:DAC

[Soh61] Louis B. Sohn. Disarmament and arms control by territories. Bulletin of the Atomic Scientists, 17(4):130–133, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sohn:1962:ZDI

[Soh62] Louis B. Sohn. Zonal disarmament and inspection: Variations on a theme. Bulletin of the Atomic Scientists, 18(7):4–7, 10, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sohn:1964:FES

[Soh64] Louis B. Sohn. Forum: European security — interrelation of political, military, and economic factors. Bulletin of the Atomic Scientists, 20(10):16–18, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sokolsky:1987:CDC

[Sok87] Joel J. Sokolsky. Canadian defense: The case for the new policy. Bulletin of the Atomic Scientists, 43(8):13–16, October 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sokolsky:1991:NWO

[Sok91] Joel J. Sokolsky. New world order: A Canadian surprise? Bulletin of the Atomic Scientists, 47(5):35, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sokolski:2001:MI

[Sok01] Henry D. Sokolski. Mission impossible. Bulletin of the Atomic Scientists, 57(2):62–68, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/2/62.

Sokolski:2016:SNQ

[Sok16] Henry Sokolski. Six nuclear questions for the next president. Bulletin of the Atomic Scientists, 72(5):278–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sokolski:2020:NTW

[Sok20] Henry Sokolski. The NPT turns 50: Will it get to 60? Bulletin of the Atomic Scientists, 76(2):63–67, 2020. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1728964.

Sokolski:2022:CUW

[Sok22] Henry Sokolski. A China–US war in space: The after-action report. Bulletin of the Atomic Scientists, 78(1):11–16, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.2014230.

Solo:1966:LES

[Sol66] Robert Solo. Letter to the Editor: Scientific magic and economic development. Bulletin of the Atomic Scientists, 22(9):25– 26, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Solomon:1968:BBV

[Sol68] Richard H. Solomon. Books: Beyond Vietnam: The United States and Asia, by Edwin O. Reischauer. Bulletin of the Atomic Scientists, 24(6):39–40, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sollie:1970:PEA

[Sol70] Finn Sollie. The political experiment in Antarctica. Bulletin of the Atomic Scientists, 26(10):16–21, December 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Antarctica since the IGY.

Solomon:1976:LNV

[Sol76] B. Solomon. Letter: Needed voicing. Bulletin of the Atomic Scientists, 32(9):7, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Solomon:1977:LN

[Sol77] Ben Solomon. Letter: NOMOR. Bulletin of the Atomic Scientists, 33(6):2–3, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel77j].

Solon:1979:CPH

[Sol79] Leonard R. Solon. Commentary: A public health approach to microwave and radiofrequency radiation. Bulletin of the Atomic Scientists, 35(8):51–55, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Solo:1980:BRB

[Sol80a] Robert A. Solo. Book review: Technology and East–West Trade. Bulletin of the Atomic Scientists, 36(8):48, 50, October 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Solo:1980:SSR

[Sol80b] Robert A. Solo. The saga of synthetic rubber. Bulletin of the Atomic Scientists, 36(4):31–36, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Solomon:1980:CRR

[Sol80c] Ben Solomon. Commentary: The right to refuse. Bulletin of the Atomic Scientists, 36(7):58, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Solon:1990:LUS

[Sol90] Leonard R. Solon. Letter: U.S. should aid Israel, nuclear weapons or no. Bulletin of the Atomic Scientists, 46(7):45–46, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei90, Gel90].

Solo:1991:NWO

[Sol91] Pam Solo. New world order: Talking law, waging war. Bulletin of the Atomic Scientists, 47(5):25–26, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mac91].

Solo:1995:TA

[Sol95] Pam Solo. Trade as aid. Bulletin of the Atomic Scientists, 51 (4):54–56, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sommer:1983:BND

[Som83a] Mark Sommer. Breaking the nuclear deadlock: Forging a preservative defense. Bulletin of the Atomic Scientists, 39(7):5–7, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Somssich:1983:CND

[Som83b] Peter Somssich. Commentary: News from Darmstadt. Bulletin of the Atomic Scientists, 39(4):47, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See followup [Som84d].

Somerville:1984:CRJ

[Som84a] John Somerville. Commentary: Reagan’s joke. Bulletin of the Atomic Scientists, 40(9):54, 56–57, November 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sommer:1984:BRB

[Som84b] Mark Sommer. Book review: Defense without the Bomb, by The Alternative Defence Commission. Bulletin of the Atomic Scientists, 40(2):41–42, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sommers:1984:BRB

[Som84c] Mark Sommers. Book review: Toward Nuclear Disarmament and Global Security: A Search for Alternatives, edited by Burns H. Weston. Bulletin of the Atomic Scientists, 40(10):30–31, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Somssich:1984:CTN

[Som84d] Peter Somssich. Commentary: Teaching nuclear war in West Germany. Bulletin of the Atomic Scientists, 40(7):50–51, August/ September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Som83b].

Sommer:1991:BRM

[Som91] Mark Sommer. Book review: Military Keynesianism: The Western Alliance after INF: Redefining U.S. Policy toward Europe and the Soviet Union, by Michael R. Lucas. Bulletin of the Atomic Scientists, 47(1):44–45, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Soman:2005:DI

[Som05] Appu K. Soman. Diplomatic illusions. Bulletin of the Atomic Scientists, 61(2):65–66, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/2/65.

Somerville:2020:FOA

[Som20] Richard C. J. Somerville. Facts and opinions about climate change. Bulletin of the Atomic Scientists, 76(6):331–335, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2020.1846422.

Sorensen:1976:WE

[Sør76] Bent Sørensen. Wind energy. Bulletin of the Atomic Scientists, 32(7):38–45, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tor77c].

Sorensen:1977:NNB

[Sør77] Bent Sørensen. No neutron bombs for us, please. Bulletin of the Atomic Scientists, 33(10):7, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Open letter to the U.S. President.

Sorensen:1980:CDA

[Sør80] Bent Sørensen. Commentary from Denmark: Anywhere but here! Bulletin of the Atomic Scientists, 36(4):55–56, April 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sorensen:1981:PRD

[Sør81] Bent Sørensen. Perspective: Redefining defense. Bulletin of the Atomic Scientists, 37(7):1, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sorensen:1995:FF

[Sør95] Georg Sørensen. Four futures. Bulletin of the Atomic Scientists, 51(4):69–72, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Soth:1967:RCW

[Sot67] Lauren Soth. Reports: Closing the world food gap. Bulletin of the Atomic Scientists, 23(9):40–42, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Soxman:1980:CTT

[Sox80] Kenneth A. Soxman. Commentary: TMI, Tet and the media. Bulletin of the Atomic Scientists, 36(3):64, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [KKHK79].

Soyfer:1989:BRB

[Soy89] Valery N. Soyfer. Book review: Science, Philosophy, and Human Behavior in the Soviet Union, by Loren R. Graham. Bulletin of the Atomic Scientists, 45(7):39–40, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Snow:1982:CMS

[SP82] Donald M. Snow and Christopher Paine. Commentary: MX smokescreen. Bulletin of the Atomic Scientists, 38(3):45–46, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pai81b].

SSBNAS:1969:BAW

[Spa69] Space Science Board, National Academy of Sciences. Beyond Apollo: Where? a prospectus. Bulletin of the Atomic Scientists, 25(7):68, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1995:CMU

[Spa95a] Lauren Spain. The C-17: a $340 million ugly duckling. Bulletin of the Atomic Scientists, 51(5):46–47, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1995:CPF

[Spa95b] Lauren Spain. Chasing the phantom fleet. Bulletin of the Atomic Scientists, 51(5):45–46, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1995:CB

[Spa95c] Lauren Spain. The competition has bowed out. Bulletin of the Atomic Scientists, 51(5):39, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1995:DMD

[Spa95d] Lauren Spain. The dream of missile defense. Bulletin of the Atomic Scientists, 51(5):49–50, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1995:SB

[Spa95e] Lauren Spain. A stealthy $72 billion. Bulletin of the Atomic Scientists, 51(5):46, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1995:TTR

[Spa95f] Lauren Spain. Treat the troops right — and save money. Bulletin of the Atomic Scientists, 51(5):42, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spain:1996:RDC

[Spa96] Lauren Spain. Report: Doomsday clock: Cool site of the day. Bulletin of the Atomic Scientists, 52(2):15–16, March/April 1996.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spedding:1949:CAA

[Spe49] F. H. Spedding. Chemical aspects of the Atomic Energy Problem. Bulletin of the Atomic Scientists, 5(2):48–50, February 1949. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Speroff:1954:LER

[Spe54] B. J. Speroff. Letter to the Editor: The remedies versus the menace. Bulletin of the Atomic Scientists, 10(8):337, October 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Comment on [Har54].

Speier:1955:WP

[Spe55] Hans Speier. War and peace. Bulletin of the Atomic Scientists, 11 (9):346–349, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SPPCSTAC:1958:EIC

[Spe58] Special Project Planning Committee on Security Through Arms Control. Establishing international control of nuclear explosions. Bulletin of the Atomic Scientists, 14(7):262–269, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sperry:1966:MBH

[Spe66] Roger W. Sperry. Mind, brain, and humanist values. Bulletin of the Atomic Scientists, 22(7):2–6, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Jor67, Fre67, vB67].

Speier:1968:BBF

[Spe68] Hans Speier. Books: The Future of Germany, by Karl Jaspers. Bulletin of the Atomic Scientists, 24(10):22–26, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spender:1972:CCV

[Spe72] Stephen Spender. A center of critical values. Bulletin of the Atomic Scientists, 28(8):2–6, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Speirs:1974:CRW

[Spe74] D. C. Speirs. Communications: Radioactive wastes on ice. Bulletin of the Atomic Scientists, 30(1):2, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, ZSA74].

Sperling:1975:LRR

[Spe75] William Sperling. Letters: Recommended reading. Bulletin of the Atomic Scientists, 31(3):2, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bur74, Dre74].

Speth:1978:NR

[Spe78] Gus Speth. The nuclear recession. Bulletin of the Atomic Scientists, 34(4):24–27, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Speth:1981:CWE

[Spe81] Gus Speth. Commentary: What to expect in the year 2000. Bulletin of the Atomic Scientists, 37(3):59–60, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1985:ONP

[Spe85a] Leonard S. Spector. 1985 outlook: Nuclear proliferation: the pace quickens. Bulletin of the Atomic Scientists, 41(1):11–14, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [RDS85].

Spector:1985:GNB

[Spe85b] Leonard S. Spector. Good news, bad news on proliferation. Bulletin of the Atomic Scientists, 41(8):16–18, September 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1986:NS

[Spe86] Leonard S. Spector. Nuclear smugglers. Bulletin of the Atomic Scientists, 42(6):34–36, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1987:NPW

[Spe87a] Leonard S. Spector. Nuclear proliferation: who’s next? Bulletin of the Atomic Scientists, 43(4):17–20, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1987:WRP

[Spe87b] Leonard S. Spector. Washington report: Pakistani smuggling riles Congress. Bulletin of the Atomic Scientists, 43(8):3–4, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1989:CSK

[Spe89a] Leonard S. Spector. Commercial satellites: Keep the skies open. Bulletin of the Atomic Scientists, 45(7):15–16, 18–20, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1989:PNP

[Spe89b] Leonard S. Spector. Proliferation: New players in the nuclear game. Bulletin of the Atomic Scientists, 45(1):29–32, January/ February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Fot89].

Speth:1989:PET

[Spe89c] James Gustave Speth. Perspective: Energy technology for survival. Bulletin of the Atomic Scientists, 45(2):9–10, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spencer:1995:OMC

[Spe95a] Metta Spencer. A one-man crusade. Bulletin of the Atomic Scientists, 51(4):66, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spencer:1995:PS

[Spe95b] Metta Spencer. “Political” scientists. Bulletin of the Atomic Scientists, 51(4):62–68, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sperling:2018:EVA

[Spe18] Daniel Sperling. Electric vehicles: Approaching the tipping point. Bulletin of the Atomic Scientists, 74(1):11–18, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1413055.

Spitzer:1949:FSC

[Spi49] Lymon Spitzer. FAS Scientists’ Committee on Loyalty Problem suggests improvement in clearance procedures. Bulletin of the Atomic Scientists, 5(11):299, November 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spitalny:1950:NYC

[Spi50] Max Spitalny. New York Committee on Atomic Information. Bulletin of the Atomic Scientists, 6(2):60, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spiegel:1954:CWC

[Spi54] John P. Spiegel. Cry wolf, cry havoc! Bulletin of the Atomic Scientists, 10(4):134–135, 144, April 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spingarn:1957:FML

[Spi57a] Jerome H. Spingarn. Five months in London. Bulletin of the Atomic Scientists, 13(7):257–261, September 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spingarn:1957:HSW

[Spi57b] Jerome H. Spingarn. The Humphrey Subcommittee: Was it worthwhile? Bulletin of the Atomic Scientists, 13(6):224–227, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spingarn:1958:DWS

[Spi58] Jerome H. Spingarn. Disarmament: The Washington scene. Bulletin of the Atomic Scientists, 14(9):385–387, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spingarn:1959:WTB

[Spi59] Jerome Spingarn. Will a test ban treaty be signed? Bulletin of the Atomic Scientists, 15(8):333–336, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1960:CSN

[Spi60] Bernard I. Spinrad. Central station nuclear reactors. Bulletin of the Atomic Scientists, 16(10):403–408, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spitzer:1961:FT

[Spi61] Lyman Spitzer, Jr. Flying telescopes. Bulletin of the Atomic Scientists, 17(5–6):191–194, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spiro:1963:LES

[Spi63] Thomas G. Spiro. Letter to the Editor: Space research. Bulletin of the Atomic Scientists, 19(5):34, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spilhaus:1964:OWW

[Spi64] Athelstan Spilhaus. Oceanography: A wet and wondrous journey. Bulletin of the Atomic Scientists, 20(10):11–15, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1965:RNE

[Spi65] Bernard I. Spinrad. Ramifications of nuclear energy. Bulletin of the Atomic Scientists, 21(5):21–24, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1966:BBN

[Spi66a] Bernard I. Spinrad. Books: In the Name of Science, by H. L. Nieburg. Bulletin of the Atomic Scientists, 22(9):22–23, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1966:FWN

[Spi66b] Bernard I. Spinrad. Forum: Why not national laboratories? Bulletin of the Atomic Scientists, 22(4):20–23, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Par66c].

Spinrad:1966:LEA

[Spi66c] Bernard I. Spinrad. Letter to the Editor: Atomic transition. Bulletin of the Atomic Scientists, 22(2):34, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Par66b].

Spinrad:1967:LEN

[Spi67] Bernard I. Spinrad. Letter to the Editor: New role for the national laboratories. Bulletin of the Atomic Scientists, 23(1):30– 31, January 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei66b].

Spinrad:1971:AEC

[Spi71a] Bernard I. Spinrad. America’s energy crisis: Reality of hysteria? Bulletin of the Atomic Scientists, 27(7):3–6, September 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dav72] and reply [Spi72b].

Spinrad:1971:IS

[Spi71b] Bernard I. Spinrad. Implications of SALT. Bulletin of the Atomic Scientists, 27(1):22–25, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1971:NEG

[Spi71c] Bernard I. Spinrad. The new emission guidelines. Bulletin of the Atomic Scientists, 27(7):7, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1972:BBS

[Spi72a] Bernard I. Spinrad. Books: The Social Organization of Electric Power Supply in Modern Societies, by Philip Sporn. Bulletin of the Atomic Scientists, 28(9):48, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1972:CR

[Spi72b] Bernard I. Spinrad. Communications: A reply. Bulletin of the Atomic Scientists, 28(3):2–3, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav72, Spi71a].

Spinrad:1972:CDC

[Spi72c] Bernard I. Spinrad. Communications: In defense of CW. Bulletin of the Atomic Scientists, 28(1):2–3, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1974:WWW

[Spi74] Bernard I. Spinrad. Where are we? On war and peace and NPT and safeguards. Bulletin of the Atomic Scientists, 30(1):34–38, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1975:BBT

[Spi75] Bernard I. Spinrad. Books: A Time to Choose: America’s Energy Future. Final Report by the Energy Policy Project of the Ford Foundation. Bulletin of the Atomic Scientists, 31(8):47–48, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spinrad:1983:NPN

[Spi83] Bernard Spinrad. Nuclear power and nuclear weapons: the connection is tenuous. Bulletin of the Atomic Scientists, 39(2):42–47, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [LLS83, Roc83b].

Spiers:1988:PG

[Spi88] Edward M. Spiers. Poison gases. Bulletin of the Atomic Scientists, 44(8):12, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sporn:1950:PIA

[Spo50] Philip Sporn. Prospects in industrial application of atomic energy. Bulletin of the Atomic Scientists, 6(10):303–306, 320, October 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sporn:1951:HCP

[Spo51] Philip Sporn. How can private industry best participate in the development of atomic energy. Bulletin of the Atomic Scientists, 7(2):48–50, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sporn:1953:PNP

[Spo53] Philip Sporn. Prospects for nuclear power. Bulletin of the Atomic Scientists, 9(4):129–140, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sponsler:1962:NST

[Spo62] George C. Sponsler. Needed: Scientists on top. Bulletin of the Atomic Scientists, 18(6):17–20, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [BE62, Sti62].

Sponsler:1964:MRS

[Spo64] George C. Sponsler. The military role in space. Bulletin of the Atomic Scientists, 20(6):31–34, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smy64].

Sponsler:1969:TT

[Spo69] George C. Sponsler. Technology for Taiwan. Bulletin of the Atomic Scientists, 25(6):31–35, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sporn:1972:CSC

[Spo72] Philip Sporn. On the control of science: a critique. Bulletin of the Atomic Scientists, 28(6):13–19, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sproull:1969:UF

[Spr69] Robert L. Sproull. Universities of the future. Bulletin of the Atomic Scientists, 25(9):43–44, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spratt:2019:RCC

[Spr19] David Spratt. Revisiting the climate collapse: The view from Nuuk in the year 2070. Bulletin of the Atomic Scientists, 75 (6):280–285, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1680050.

Spratt:2025:SRH

[Spr25] David Spratt. Is scientific reticence hindering climate understanding? Bulletin of the Atomic Scientists, 81(2):107–113, 2025. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2025.2464442.

Spurgeon:1972:NAF

[Spu72] David Spurgeon. A new approach to foreign aid: The IDRC of Canada. Bulletin of the Atomic Scientists, 28(9):33–36, November 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spurgeon:1973:TFS

[Spu73] David Spurgeon. Tightening food supplies: The nutrition crunch: A world view. Bulletin of the Atomic Scientists, 29(8):50–54, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spurgeon:1980:CNA

[Spu80] David Spurgeon. Commentary from Nairobi: Agroforestry. Bulletin of the Atomic Scientists, 36(1):58–61, January 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Squires:1975:LTN

[Squ75] Vera Moore Squires. Letter: Toward a new era. Bulletin of the Atomic Scientists, 31(10):2, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75f].

Squires:1977:LVP

[Squ77] Vera Moore Squires. Letter: Voices from the past. Bulletin of the Atomic Scientists, 33(2):2, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel76d].

Squassoni:2012:HCC

[Squ12] Sharon Squassoni. How can countries ensure that the nuclear security summit does not lose momentum and become just another gathering?: Position: Set meaningful goals to make the summit count. Bulletin of the Atomic Scientists, 68(2):78–80, March/ April 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/68/2/78.

Squassoni:2013:LNS

[Squ13] Sharon Squassoni. The limited national security implications of civilian nuclear decline. Bulletin of the Atomic Scientists, 69(2): 22–33, March/April 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/2/22.

Squassoni:2016:NG

[Squ16] Sharon Squassoni. The nuclear Google. Bulletin of the Atomic Scientists, 72(5):282–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Squassoni:2017:ISN

[Squ17a] Sharon Squassoni. The incredible shrinking nuclear offset to climate change. Bulletin of the Atomic Scientists, 73(1):17–26, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2016.1264208.

Squassoni:2017:TFL

[Squ17b] Sharon Squassoni. Through a fractured looking-glass: Trump’s nuclear decisions so far. Bulletin of the Atomic Scientists, 73 (6):370–375, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1388667.

Squassoni:2021:NED

[Squ21a] Sharon Squassoni. Nuclear energy: A distraction on the road to climate solutions. Bulletin of the Atomic Scientists, 77(6): 290–294, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1990488.

Squassoni:2021:WBS

[Squ21b] Sharon Squassoni. Why Biden should abandon the great power competition narrative. Bulletin of the Atomic Scientists, 77(1): 3–4, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/00963402.2020.1859854.

Sabato:1980:ATW

[SR80] Jorge A. Sabato and Jairam Ramesh. Atoms for the Third World. Bulletin of the Atomic Scientists, 36(3):36–43, March 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Shoenholz:1981:CHP

[SR81a] Stephen Shoenholz and Joel Ray. Commentary: High power lines. Bulletin of the Atomic Scientists, 37(2):59–61, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ray80].

Solon:1981:CRR

[SR81b] Leonard R. Solon and Kenneth Rosenberg. Commentary: The release of radioiodines in a nuclear emergency. Bulletin of the Atomic Scientists, 37(8):54–56, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salzberg:1967:FS

[SS67] H. W. Salzberg and Herbert I. Schiller. Forum: On Schiller. Bulletin of the Atomic Scientists, 23(6):63–64, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sch67b].

Schapsmeier:1969:COY

[SS69] Edward L. Schapsmeier and Frederick H. Schapsmeier. Communications: The old and the young. Bulletin of the Atomic Scientists, 25(8):48, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab69k, Ame69].

Smith:1973:HAP

[SS73] J. V. Smith and I. M. Steele. How the Apollo Program changed the geology of the Moon: Suggestions for future studies. Bulletin of the Atomic Scientists, 29(9):11–15, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1976:NED

[SS76] Philip B. Smith and Ruud Spanhoff. The nuclear energy debate in The Netherlands. Bulletin of the Atomic Scientists, 32(2):41– 44, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi76].

Sorensen:1980:CDR

[SS80] Bent Sørensen and Vaclav Smil. Commentary from Denmark: On renewable energies. Bulletin of the Atomic Scientists, 36(7):55– 57, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi79].

Spector:1988:CAR

[SS88] Leonard S. Spector and Shelley A. Stahl. Cooling the arms race in South Asia. Bulletin of the Atomic Scientists, 44(3):32–33, April 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schmidt:1989:EGN

[SS89] Max Schmidt and Wolfgang Schwarz. From East Germany: It’s NATO’s move now. Bulletin of the Atomic Scientists, 45(7):5– 6, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Møl90].

Spector:1990:NRT

[SS90a] Leonard S. Spector and Jacqueline R. Smith. NPT review: Treaty review: Deadlock damages nonproliferation. Bulletin of the Atomic Scientists, 46(10):39–44, December 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Spector:1990:NAS

[SS90b] Leonard S. Spector and Jacqueline R. Smith. Nuclear Ambitions: the Spread of Nuclear Weapons, 1989–1990. Westview Press, Boulder, CO, USA, 1990. ISBN 0-8133-8075-8, 0-8133-8074-X. 450 pp. LCCN U264 .S629.

Siegel:2005:WMP

[SS05] Jonas Siegel and Josh Schollmeyer. W’s multilateralism primer. Bulletin of the Atomic Scientists, 61(2):36–37, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/2/36.

Schwellenbach:2006:NL

[SS06] Nick Schwellenbach and Peter D. H. Stockton. Nuclear lockdown. Bulletin of the Atomic Scientists, 62(6):44–50, November/ December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/62/
6/44.

Sinclair:2024:ANM

[SS24] Samuel Justin Sinclair and David A. Silbersweig. Apocalypse now? Mortality and mental health correlates of the Doomsday Clock. Bulletin of the Atomic Scientists, 81(2):126–134, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2439762.

Sicular:1968:NWE

[SSC68] George M. Sicular, W. R. Derrick Sewell, and James A. Crutchfield. Nawapa: A wet exchange. Bulletin of the Atomic Scientists, 24(7):35–37, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lew65b].

Samuel:2010:BFC

[SSK10] Gabrielle Samuel, Michael J. Selgelid, and Ian Kerridge. Back to the future: Controlling synthetic life science trade in DNA sequences. Bulletin of the Atomic Scientists, 66(5):9–20, September/October 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/66/5/9.

Stratton:1983:CNF

[SSR83] R. W. Stratton, Peter Schenkel, and Earl C. Ravenal. Commentary: No first use: a view from the United States. Bulletin of the Atomic Scientists, 39(7):60–62, August/September 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rav83].

Sax:1955:RHH

[SSW^{+}55] Karl Sax, B. P. Sonnenblick, Sewall Wright, Joshua Lederberg, and C. H. Waddington. Radiation and human heredity: Comment from geneticists. Bulletin of the Atomic Scientists, 11(10):364– 366, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strozier:2022:PPN

[ST22] Charles B. Strozier and David M. Terman. Putin’s psychology and nuclear weapons: The fundamentalist mindset. Bulletin of the Atomic Scientists, 78(6):310–314, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2132736.

Star:1948:LIP

[Sta48] Shirley A. Star. Loyalty investigations — a poll of atomic scientists. Bulletin of the Atomic Scientists, 4(7):218, July 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stace:1949:SMH

[Sta49] Walter T. Stace. Science, materialism, and the human spirit: The need for a secular ethic. Bulletin of the Atomic Scientists, 5(6– 7):197–198, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Staley:1958:TMS

[Sta58] Eugene Staley. Technology: Master or servant? Bulletin of the Atomic Scientists, 14(10):417–420, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stadler:1960:LED

[Sta60a] Anne M. Stadler. Letter to the Editor: Deeds, not words. Bulletin of the Atomic Scientists, 16(10):374, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab60a].

Staley:1960:SDF

[Sta60b] Eugene Staley. Scientific developments and foreign policy. Bulletin of the Atomic Scientists, 16(1):7–13, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stadler:1961:LER

[Sta61] Anne M. Stadler. Letter to the Editor: The real question. Bulletin of the Atomic Scientists, 17(3):121, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stadler:1962:VTU

[Sta62a] Anne M. Stadler. Voluntary ‘tax’ for U.N. Bulletin of the Atomic Scientists, 18(1):38, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Staley:1962:RU

[Sta62b] Eugene Staley. Revenue for the U.N. Bulletin of the Atomic Scientists, 18(7):43–44, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stanley:1963:FGM

[Sta63] Timothy W. Stanley. Forum: The gall of Monsieur Gallois. Bulletin of the Atomic Scientists, 19(8):27–30, October 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gal63].

Stanford:1964:BBS

[Sta64] George S. Stanford. Books: Strategy for Survival, by Thomas L. Martin, Jr., and Donald C. Latham. Bulletin of the Atomic Scientists, 20(4):35–36, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SBSG:1971:DI

[Sta71] Stanford Biology Study Group. The destruction of Indochina. Bulletin of the Atomic Scientists, 27(5):36–40, May 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Starr:1973:REC

[Sta73] Chauncey Starr. Realities of the energy crisis. Bulletin of the Atomic Scientists, 29(7):15–20, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stang:1979:PNP

[Sta79] Fredrik Stang. Perspective: Nobel Peace Laureates: Carl von Ossietzky, 1935. Bulletin of the Atomic Scientists, 35(2):1, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Von Ossietzky was awarded the Prize, but the Nazi government refused to release him from a concentration camp, and decreed that no German in the future could accept any Nobel Prize. Von Ossietzky died in hospital under Nazi surveillance in May 1938.

Stanford:1980:CLG

[Sta80] George S. Stanford. Commentary: A legalistic game? Bulletin of the Atomic Scientists, 36(9):60, November 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [STRS80].

Stanley:1981:PND

[Sta81] C. Maxwell Stanley. Perspective: New definition for national security. Bulletin of the Atomic Scientists, 37(3):1, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Staar:1985:BRB

[Sta85] Richard F. Staar. Book review: The Warsaw Pact: Alliance in Transition?, edited by David Holloway and Jane M. O. Sharp. Bulletin of the Atomic Scientists, 41(5):42–43, May 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sha85b].

Stauffer:1987:LCV

[Sta87] James Stauffer. Letter: Calculations vs. principles. Bulletin of the Atomic Scientists, 43(9):51–52, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [SHT^{+}87].

Stanbro:1993:LFA

[Sta93] Helen Stanbro. Letter: False alarm. Bulletin of the Atomic Scientists, 49(4):47, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano93-27].

Stampfli:1996:LA

[Sta96] John F. Stampfli. Letter: 14 agencies. Bulletin of the Atomic Scientists, 52(6):3, 63, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pra96].

Stanford:1997:LCM

[Sta97] George S. Stanford. Letter: Chernobyl, moment by moment. Bulletin of the Atomic Scientists, 53(3):3, 59–60, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hes97, Sic96a] and comments [HS97].

Stacy:2000:PPH

[Sta00] Susan M. Stacy. Proving the principle: a history of the Idaho National Engineering and Environmental Laboratory, 1949–1999. Idaho Operations Office of the Department of Energy, Idaho Falls, ID, USA, 2000. ISBN 0-16-059185-6. xiv + 326 pp. LCCN

TK9202.S73 P76 2000. URL http://purl.access.gpo.gov/
GPO/LPS15610.

Speth:1974:PRF

[STC74] J. Gustave Speth, Arthur R. Tamplin, and Thomas B. Cochran. Plutonium recycle: The fateful step: Impending move to reprocess fuel would escalate the risks of nuclear power. Bulletin of the Atomic Scientists, 30(9):14–22, November 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MC75].

Stern:1946:MSR

[Ste46] Otto Stern. Military support of research vs. national science legislation. Bulletin of the Atomic Scientists, 2(9–10):31, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sterenberg:1952:UNA

[Ste52] Elizabeth Sterenberg. United Nations atomic energy news: Disarmament news. Bulletin of the Atomic Scientists, 8(4):116–118, 126, April 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stevenson:1953:STL

[Ste53] Adlai E. Stevenson. Some thoughts on loyalty. Bulletin of the Atomic Scientists, 9(1):16, 28, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stead:1957:SFP

[Ste57a] William H. Stead. The sun and foreign policy. Bulletin of the Atomic Scientists, 13(3):86–90, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinbicker:1957:SE

[Ste57b] Paul G. Steinbicker. Shelter or evacuation? Bulletin of the Atomic Scientists, 13(5):166–168, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stevenson:1958:NSN

[Ste58] Adlai Stevenson. New steps needed. Bulletin of the Atomic Scientists, 14(7):255, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stehn:1960:PPP

[Ste60a] John R. Stehn. Project Plowshare: Peaceful uses of nuclear explosives. Bulletin of the Atomic Scientists, 16(10):416, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Slightly modified version of article in Mase Newsletter, 28 May 1960.

Stern:1960:RCI

[Ste60b] Frederick Martin Stern. Readers’ comment: Incomplete deterrence. Bulletin of the Atomic Scientists, 16(10):414–415, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stevenson:1960:SRU

[Ste60c] Adlai Stevenson. [Shock at resumption of underground testing]. Bulletin of the Atomic Scientists, 16(3):85, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stewart:1961:SSC

[Ste61] Bruce Stewart. Science and social change. Bulletin of the Atomic Scientists, 17(7):267–270, 286, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinbach:1962:SPP

[Ste62] H. Burr Steinbach. Scientists and public policy. Bulletin of the Atomic Scientists, 18(3):11–13, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stevenson:1965:FSS

[Ste65a] Thomas H. Stevenson. Forum: State and science five centuries ago. Bulletin of the Atomic Scientists, 21(1):28–29, January 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stewart:1965:LAS

[Ste65b] Bruce Stewart. Letters: Art, science, and …. Bulletin of the Atomic Scientists, 21(2):34–35, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stegemen:1969:SA

[Ste69a] Beatrice Stegemen. Science as art. Bulletin of the Atomic Scientists, 25(4):27–30, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sternberg:1969:ACE

[Ste69b] Sidney Sternberg. Automatic checkout equipment — the Apollo Hippocrates. Bulletin of the Atomic Scientists, 25(7):84–87, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sternglass:1969:CRa

[Ste69c] Ernest J. Sternglass. Comments: A reply. Bulletin of the Atomic Scientists, 25(8):29–32, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e, Sag69, Wei69b, Wor69, Tam69, NW70, Ste70].

Sternglass:1969:CRb

[Ste69d] Ernest J. Sternglass. Comments: A reply. Bulletin of the Atomic Scientists, 25(10):29–34, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e, Tam69].

Sternglass:1969:IMN

[Ste69e] Ernest J. Sternglass. Infant mortality and nuclear tests. Bulletin of the Atomic Scientists, 25(4):18–20, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Sag69, Sag69, Wei69b, Wor69, Ano70b, Zel70] and reply [Ste69c].

Sternglass:1969:MAM

[Ste69f] Ernest J. Sternglass. Missiles and anti-missiles … six views: Can the infants survive? Bulletin of the Atomic Scientists, 25 (6):26–27, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sternglass:1970:CR

[Ste70] Ernest J. Sternglass. Communications: A reply. Bulletin of the Atomic Scientists, 26(5):41–42, 47, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NW70].

Sternglass:1972:CRR

[Ste72a] Ernest J. Sternglass. Communications: Radiation risks. Bulletin of the Atomic Scientists, 28(6):4–5, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Led72, Led72].

Stewart:1972:CSE

[Ste72b] Bruce Stewart. Communications: Space exploits. Bulletin of the Atomic Scientists, 28(10):3, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lev73].

Sternig:1973:CSS

[Ste73a] John Sternig. Communications: Solar storage. Bulletin of the Atomic Scientists, 29(4):56, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stewart:1973:IO

[Ste73b] Harris B. Stewart, Jr. The international ocean. Bulletin of the Atomic Scientists, 29(8):38, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stecker:1974:RAB

[Ste74] Floyd W. Stecker. The role of antimatter in Big-Bang cosmology. Bulletin of the Atomic Scientists, 30(1):12–16, January 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano74e].

Steiner:1975:BAS

[Ste75a] A. Steiner. Baptism of atomic scientists. Bulletin of the Atomic Scientists, 31(2):21–28, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stern:1975:WED

[Ste75b] Robert J. Stern. When the experts disagree: The public needs a better system to separate fact from opinion. Bulletin of the Atomic Scientists, 31(2):29–31, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [For75].

Steiner:1976:BST

[Ste76a] Arthur Steiner. Books: Sovereignty and technology: Technology and International Politics: The Crisis of Wishing, by William R. Kintner and Harvey Sicherman; The Nuclear Age; Nuclear Proliferation Problems; Safeguards Against Nuclear Proliferation. Bulletin of the Atomic Scientists, 32(2):65–66, February 1976. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinitz:1976:LC

[Ste76b] Willis S. Steinitz. Letter: A catalyst. Bulletin of the Atomic Scientists, 32(9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Steinzor:1976:LDM

[Ste76c] Bernard Steinzor. Letter: A disturbing matter. Bulletin of the Atomic Scientists, 32(9):4, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bru76].

Steinberg:1977:LAB

[Ste77a] Meyer Steinberg. Letter: The accelerator–breeder. Bulletin of the Atomic Scientists, 33(10):2–3, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pen77a].

Steiner:1977:BEA

[Ste77b] Arthur Steiner. Books: The enormity of the arms race: Armaments and Disarmament in the Nuclear Age: a Handbook. Bulletin of the Atomic Scientists, 33(8):63–64, October 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [You78b].

Stewart:1977:SNE

[Ste77c] Bruce Stewart. Some nuclear explosions will be necessary: atomic ‘shock treatment’ may be the only cure for man’s habit of uncontrolled arms proliferation. Bulletin of the Atomic Scientists, 33 (8):51–54, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stevenson:1978:LCB

[Ste78] Kalmar R. Stevenson. Letter: Continuing body dount. Bulletin of the Atomic Scientists, 34(5):5, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bar77b, Nat77].

Steiner:1980:BRB

[Ste80a] Arthur Steiner. Book review: The Nuclear Question: The United States and Nuclear Weapons. Bulletin of the Atomic Scientists, 36(7):42–45, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steiner:1980:CKP

[Ste80b] Oscar H. Steiner. Commentary: Knowledge is power. Bulletin of the Atomic Scientists, 36(4):65, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steiner:1981:REH

[Ste81a] Arthur Steiner. 1980 Rabinowitch Essay on how to eliminate the threat of nuclear war: Openness in a nuclear world. Bulletin of the Atomic Scientists, 37(3):39–44, March 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stephenson:1981:CWH

[Ste81b] Thomas A. Stephenson. Commentary: Working for human rights. Bulletin of the Atomic Scientists, 37(7):54–56, August/ September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste82b].

Stein:1982:UUH

[Ste82a] David Stein. US/USSR hostages for peace. Bulletin of the Atomic Scientists, 38(2):48, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stephenson:1982:CWH

[Ste82b] Thomas A. Stephenson. Commentary: Working for human rights. Bulletin of the Atomic Scientists, 38(3):49, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste81b].

Stein:1983:EPU

[Ste83a] Daniel L. Stein. Electromagnetic pulse — the uncertain certainty. Bulletin of the Atomic Scientists, 39(3):52–56, March 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinberg:1983:TME

[Ste83b] Gerald M. Steinberg. Two missiles in every garage. Bulletin of the Atomic Scientists, 39(8):43–48, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sha84a].

Stewart:1984:S

[Ste84] Bruce Stewart. Out of service. Bulletin of the Atomic Scientists, 40(6):52–53, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [But84].

Stein:1985:LDB

[Ste85a] Jonathan B. Stein. Letter: Defense, build-down and MX. Bulletin of the Atomic Scientists, 41(2):52–53, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stein:1985:PPT

[Ste85b] Jonathan B. Stein. Political push, technological pull. Bulletin of the Atomic Scientists, 41(6):40–41, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinbruner:1985:USS

[Ste85c] John Steinbruner. U.S. and Soviet security perspectives. Bulletin of the Atomic Scientists, 41(7):89–93, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stein:1986:NTM

[Ste86a] Josephine Anne Stein. Nuclear tests mean new weapons. Bulletin of the Atomic Scientists, 42(9):8–11, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinberg:1986:LSR

[Ste86b] Meyer Steinberg. Letter: Safe reactors. Bulletin of the Atomic Scientists, 42(9):56–57, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei86a].

Sterner:1986:BRB

[Ste86c] Michael Sterner. Book review: Arms and Oil: U.S. Military Strategy and the Persian Gulf, by Thoms L. McNaugher. Bulletin of the Atomic Scientists, 42(3):53–54, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sternglass:1987:LGB

[Ste87] Ernest J. Sternglass. Letter: Ground-based SDI. Bulletin of the Atomic Scientists, 43(2):62–63, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steele:1988:MWH

[Ste88] Karen Dorn Steele. Making warheads: Hanford’s bitter legacy. Bulletin of the Atomic Scientists, 44(1):17–23, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steele:1989:BNS

[Ste89a] Karen Dorn Steele. Bulletin: National security ever green. Bulletin of the Atomic Scientists, 45(10):6, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steele:1989:HAN

[Ste89b] Karen Dorn Steele. Hanford: America’s nuclear graveyard. Bulletin of the Atomic Scientists, 45(8):14–23, October 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sto90].

Steele:1989:LRF

[Ste89c] Karen Dorn Steele. Last rites for first commercial reactor. Bulletin of the Atomic Scientists, 45(8):17, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steele:1989:WDB

[Ste89d] Karen Dorn Steele. What to do with B Reactor? Bulletin of the Atomic Scientists, 45(8):21, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stein:1989:BRB

[Ste89e] Jonathan Stein. Book review: Technology and War, by Martin van Creveld. Bulletin of the Atomic Scientists, 45(6):47–48, July/ August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinberg:1989:LKS

[Ste89f] Lawrence William Steinberg. Letter: Knees still bent. Bulletin of the Atomic Scientists, 45(4):55, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei89a].

Steele:1990:RHT

[Ste90a] Karen Dorn Steele. Radiation and health: Tracking down Hanford’s victims. Bulletin of the Atomic Scientists, 46(8):7–8, 46–47, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stewart:1990:RHL

[Ste90b] Alice M. Stewart. Radiation and health: Low-level radiation: the cancer controversy. Bulletin of the Atomic Scientists, 46(7):15– 18, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steel:1991:BRF

[Ste91a] Ronald Steel. Book review: Full circle: From the Finland Station: The Graying of Revolution in the Twentieth Century, by Theodore S. Hamerow. Bulletin of the Atomic Scientists, 47(3): 45, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steele:1991:RNC

[Ste91b] Karen Dorn Steele. Report: Nuclear cleanup: Hanford in hot water. Bulletin of the Atomic Scientists, 47(4):7–8, May 1991. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinbruner:1991:NWO

[Ste91c] John Steinbruner. New world order: The rule of law. Bulletin of the Atomic Scientists, 47(5):20, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sternman:1995:BRF

[Ste95] Mark S. Sternman. Book review: Freezing jet fighters: The Arms Production Dilemma: Contraction and Restraint in the World Combat Aircraft Industry, edited by Randall Forsberg. Bulletin of the Atomic Scientists, 51(3):67, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Steinberger:1997:HKH

[Ste97] Michael Steinberger. For Hong Kong, the horn blows at midnight. Bulletin of the Atomic Scientists, 53(1):26–28, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stern:2001:MM

[Ste01] Jessica Stern. Meeting with the muj. Bulletin of the Atomic Scientists, 57(1):42–50, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/1/42.

Stebbins:2007:AS

[Ste07a] Michael Stebbins. Agents of suspicion. Bulletin of the Atomic Scientists, 63(3):26, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/3/26.

Stebbins:2007:SS

[Ste07b] Michael Stebbins. Stumping on security. Bulletin of the Atomic Scientists, 63(6):19, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/6/19.

Steinbruner:2007:BP

[Ste07c] John Steinbruner. Burdens of proof. Bulletin of the Atomic Scientists, 63(1):59–61, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/59.

Steinbruner:2008:CS

[Ste08] John Steinbruner. Consensual security. Bulletin of the Atomic Scientists, 64(1):23–27, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/64/1/23.

Schneider:2009:WNI

[STFK09] Mycle Schneider, Steve Thomas, Antony Froggatt, and Doug Koplow. 2009 world nuclear industry status report. Bulletin of the Atomic Scientists, 65(6):1–19, November 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/65/6/1.

Stimson:1947:DUA

[Sti47] Henry L. Stimson. The decision to use the atomic bomb. Bulletin of the Atomic Scientists, 3(2):37–41, 66–67, February 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stimson:1948:ABP

[Sti48] Henry L. Stimson. The atomic bomb and peace with Russia. Bulletin of the Atomic Scientists, 4(8):237–244, August 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stirrett:1952:LEB

[Sti52] J. R. Stirrett. Letter to the Editor: Atomic Diplomacy. Bulletin of the Atomic Scientists, 8(2):65, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stirrett:1957:LEN

[Sti57] J. R. Stirrett. Letter to the Editor: A New Year wish. Bulletin of the Atomic Scientists, 13(3):113, March 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Still:1960:BPF

[Sti60] Joseph W. Still. Biology, psychology, and foreign policy. Bulletin of the Atomic Scientists, 16(1):14–17, January 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Still:1961:TVB

[Sti61] Joseph W. Still. Two views on Berlin. Bulletin of the Atomic Scientists, 17(7):294, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lan61b].

Still:1962:LES

[Sti62] Joseph W. Still. Letter to the Editor: Scientists on top. Bulletin of the Atomic Scientists, 18(9):32, November 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spo62].

Still:1963:FZP

[Sti63] Joseph W. Still. Forum: A zone of peace. Bulletin of the Atomic Scientists, 19(7):29–30, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Still:1967:CLY

[Sti67] Joseph W. Still. Comments: Love your enemies. Bulletin of the Atomic Scientists, 23(3):30, March 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Open letter to President Lyndon B. Johnson.

Still:1968:NMC

[Sti68] Joseph W. Still. Noninterference must come first. Bulletin of the Atomic Scientists, 24(8):45–47, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stivers:1978:LPD

[Sti78] William A. Stivers. Letter: Perpetuating dependence. Bulletin of the Atomic Scientists, 34(2):5–6, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mei77b].

Stillman:1983:BRB

[Sti83] Peter G. Stillman. Book reviews: Crisis Contained: The Department of Energy at Three Mile Island, by Philip L. Cantelon and Robert C. Williams; Three Mile Island: Thirty Minutes to Meltdown, by Daniel F. Ford; The Warning, by Mike Gray and Ira Rosen; Three Mile Island: Turning Point, by Bill Keisling; Accident at Three Mile Island: The Human Dimensions, edited by David L. Sills, C. P. Wolf, and Vivian B. Shelanski; Three Mile Island, by Mark Stephens. Bulletin of the Atomic Scientists, 39 (9):42–45, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stiles:2018:SLR

[Sti18] William A. Stiles Jr. Sea level rise — from my front porch. Bulletin of the Atomic Scientists, 74(2):81–90, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2018.1436807.

Stokley:1953:CDB

[Sto53] Robert W. Stokley. Civil defense and the budget. Bulletin of the Atomic Scientists, 9(7):256–258, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1955:BRBa

[Sto55a] Ursula B. Stone. Book review: Past and Future, by William T. McNeill. Bulletin of the Atomic Scientists, 11(3):94–95, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1955:BRBc

[Sto55b] Ursula Batchelder Stone. Book review: Faith and Freedom, by Barbara Ward. Bulletin of the Atomic Scientists, 11(10):376, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1955:BRBb

[Sto55c] Ursula Batchelder Stone. Book review: Speak Truth to Power. Bulletin of the Atomic Scientists, 11(9):349, November 1955. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1956:BBO

[Sto56a] Ursula Batchelder Stone. Books: The Only Way Out, by Jacob Rosen. Bulletin of the Atomic Scientists, 12(3):93, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1956:BBS

[Sto56b] Ursula Batchelder Stone. Books: Towards a Science of Peace, by Theodore F. Lentz; The Price of Peace, by Charles G. Bolté; The New Dimensions of Peace, by Chester Bowles. Bulletin of the Atomic Scientists, 12(6):231, June 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1958:BBP

[Sto58] Ursula Batchelder Stone. Books: Some Problems of the Atomic Age, by Charles Alfred Coulson; Will the Atom Unite the World?, by Angelos Angelopoulos; A United Nations Peace Force, by William R. Frye. Bulletin of the Atomic Scientists, 14(8):336, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1959:BBS

[Sto59a] Ursula Batchelder Stone. Books: The Sleepwalkers: A History of Man’s Changing Universe, by Arthur Koestler. Bulletin of the Atomic Scientists, 15(8):349–350, October 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1959:BBW

[Sto59b] Ursula Batchelder Stone. Books: World Without War, by J. D. Bernal; Rededication to Freedom, by Benjamin Ginzburg; Organizing for Peace in the Nuclear Age, edited by Arthur N. Holcombe; Freedom in a Federal World, by Everett Lee Millard; Permanent Peace, by Tom Slick; The Prerequisites for Peace, by Norman Thomas. Bulletin of the Atomic Scientists, 15(10): 427, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stowens:1960:LEW

[Sto60] Daniel Stowens. Letter to the Editor: Weapons are everywhere. Bulletin of the Atomic Scientists, 16(7):305, September 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stout:1961:RCC

[Sto61] Perry R. Stout. Reflections and comments: Controversy on Kahn. Bulletin of the Atomic Scientists, 17(7):293, September 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar61d, Kea61, Pet61, Mar61c].

Stonier:1962:LES

[Sto62] Tom Stonier. Letter to the Editor: Strategy. Bulletin of the Atomic Scientists, 18(10):31–32, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1964:ARD

[Sto64] Jeremy J. Stone. Arms race or disarmament. Bulletin of the Atomic Scientists, 20(7):20–24, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [LS65].

Stone:1965:FCA

[Sto65a] Jeremy J. Stone. Forum: Containing the arms race. Bulletin of the Atomic Scientists, 21(7):18–21, September 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1965:FPW

[Sto65b] Jeremy J. Stone. Forum: On proliferation: Where’s the danger? Bulletin of the Atomic Scientists, 21(9):15–18, November 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [TK66].

Stone:1966:ANM

[Sto66a] Jeremy J. Stone. ABM — the next MLF? Bulletin of the Atomic Scientists, 22(7):20–21, September 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1966:RMS

[Sto66b] Jeremy J. Stone. Reports: The McNamara story continues. Bulletin of the Atomic Scientists, 22(4):39–42, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1967:BNR

[Sto67] Jeremy J. Stone. Beginning of the next round? Bulletin of the Atomic Scientists, 23(10):20–25, December 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1968:BBA

[Sto68a] Marshall H. Stone. Books: The Autobiography of Bertrand Russell, 1872–1914. Bulletin of the Atomic Scientists, 24(1):54–55, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LB68].

Stout:1968:PKF

[Sto68b] Perry Stout. Power: The key to food sufficiency in India? Bulletin of the Atomic Scientists, 24(9):26–28, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stonier:1972:ISE

[Sto72] Tom Stonier. An international solar energy development decade: A proposal for global cooperation. Bulletin of the Atomic Scientists, 28(5):31–34, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1976:LRF

[Sto76a] Jeremy J. Stone. Letter: Restricting first use. Bulletin of the Atomic Scientists, 32(4):4, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kis76b].

Stopal:1976:LPP

[Sto76b] Simon A. Stopal. Letter: Playing with pure ‘Barnum’. Bulletin of the Atomic Scientists, 32(10):2, 4, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sal76].

Storrs:1977:LED

[Sto77] C. L. Storrs. Letter: Energy delusion. Bulletin of the Atomic Scientists, 33(2):3, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1978:LSE

[Sto78a] Jeremy Stone. Linking SALT to Ethiopia or unlinking it from détente. Bulletin of the Atomic Scientists, 34(6):38–39, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stoner:1978:LSB

[Sto78b] A. Douglas Stoner. Letter: Some bad, some good. Bulletin of the Atomic Scientists, 34(5):6, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha78b, Ros78b].

Stone:1979:CWS

[Sto79] Jeremy J. Stone. Commentary: Are we seeking disarmament or détente? Bulletin of the Atomic Scientists, 35(8):8–9, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stone:1985:BRB

[Sto85] Jeremy J. Stone. Book review: Deadly gambits, by Strobe Talbott. Bulletin of the Atomic Scientists, 41(1):47–48, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stoffels:1990:LGT

[Sto90] Jim Stoffels. Letter: Guided tours through plutonium reactor. Bulletin of the Atomic Scientists, 46(6):46, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste89b] and comment [Hyd91].

Stone:1996:LTM

[Sto96] Jeremy J. Stone. Less than meets the eye. Bulletin of the Atomic Scientists, 52(5):43–45, September/October 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stoffels:1997:LNS

[Sto97] Jim Stoffels. Letter: Next step: Ban tritium. Bulletin of the Atomic Scientists, 53(4):58–59, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rot97k].

Stober:1999:NSS

[Sto99] Dan Stober. Nuclear secrets: Steal this! Bulletin of the Atomic Scientists, 55(4):14–16, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/4/14.

Stober:2003:NEN

[Sto03] Dan Stober. No experience necessary. Bulletin of the Atomic Scientists, 59(2):58–63, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/59/2/58.

Stover:2016:SSU

[Sto16] Dawn Stover. Susan Southard: Unveiling the aftermath of nuclear war. Bulletin of the Atomic Scientists, 72(4):199–??, 2016. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stover:2017:BIC

[Sto17a] Dawn Stover. Bob Inglis: A conservative for climate action. Bulletin of the Atomic Scientists, 73(6):364–367, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1388661.

Stover:2017:KEC

[Sto17b] Dawn Stover. Kerry Emanuel: A climate scientist for nuclear energy. Bulletin of the Atomic Scientists, 73(1):7–12, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2016.1264205.

Stover:2017:PHG

[Sto17c] Dawn Stover. Paul Hawken: “Game on” for global warming. Bulletin of the Atomic Scientists, 73(3):145–149, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1314995.

Stover:2018:ABJ

[Sto18a] Dawn Stover. Ambassador Bonnie Jenkins: Reducing threats, increasing diversity. Bulletin of the Atomic Scientists, 74(6): 404–408, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1533218.

Stover:2018:ECE

[Sto18b] Dawn Stover. Eileen Claussen: Engaging businesses on climate action. Bulletin of the Atomic Scientists, 74(1):36–40, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1413061.

Stover:2018:GGF

[Sto18c] Dawn Stover. Garlin Gilchrist: Fighting fake news and the information apocalypse. Bulletin of the Atomic Scientists, 74 (4):283–288, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2018.1486618.

Stover:2019:DWS

[Sto19a] Dawn Stover. Dominic Woolf: Studying soil and biochar for carbon dioxide removal. Bulletin of the Atomic Scientists, 75 (5):229–235, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1654265.

Stover:2019:ECA

[Sto19b] Dawn Stover. Evangelicals for climate action. Bulletin of the Atomic Scientists, 75(2):66–72, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2019.
1580878.

Stover:2019:GJO

[Sto19c] Dawn Stover. Greg Jaczko, the outsider who ran the Nuclear Regulatory Commission. Bulletin of the Atomic Scientists, 75 (3):116–121, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1604907.

Stover:2019:MSC

[Sto19d] Dawn Stover. Marshall Shepherd: Connecting atmospheric science and society. Bulletin of the Atomic Scientists, 75(4):205– 209, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1628517.

Stover:2023:WAM

[Sto23] Dawn Stover. Will AI make us crazy? Bulletin of the Atomic Scientists, 79(5):299–303, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2245247.

Stover:2024:ILN

[Sto24] Dawn Stover. Interview: Lawrence Norden on US election security. Bulletin of the Atomic Scientists, 80(5):321–325, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2024.2388469.

Strauss:1954:HBI

[Str54a] Lewis L. Strauss. The H-bomb and the international pool proposal. Bulletin of the Atomic Scientists, 10(6):229–230, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from a speech by the author to the Los Angeles World Affairs Council on 19 April 1954.

Strauss:1954:HBW

[Str54b] Lewis L. Strauss. The H-bomb and world opinion: Chairman Strauss’ statement on Pacific tests. Bulletin of the Atomic Scientists, 10(5):163–165, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strauss:1956:EPG

[Str56] Lewis L. Strauss. [Eniwetok Proving Grounds in the Marshall Islands and my report to the President]. Bulletin of the Atomic Scientists, 12(7):263, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strauss:1958:UAE

[Str58] Lewis L. Strauss. The U.S. Atomic Energy Program: 1953– 58. Bulletin of the Atomic Scientists, 14(7):256–258, September 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strauss:1962:TDS

[Str62] Lewis L. Strauss. [The two decades since December 2, 1942]. Bulletin of the Atomic Scientists, 18(10):16, December 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strassmann:1967:BBE

[Str67] W. Paul Strassmann. Books: Modern Economic Growth: Rate, Structure, Spread, by Simon Kuznets. Bulletin of the Atomic Scientists, 23(6):52–53, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strauss:1971:SEI

[Str71] Herbert L. Strauss. Science education in India. Bulletin of the Atomic Scientists, 27(3):10–13, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strope:1972:CAR

[Str72] Walmer E. Strope. Communications: Arms race. Bulletin of the Atomic Scientists, 28(4):3, April 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pan71, Kra72].

Strohm:1978:LFN

[Str78] Holger Strohm. Letter: ‘friends’ needed. Bulletin of the Atomic Scientists, 34(4):4, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Swe77].

Street:1981:PIS

[Str81] James H. Street. Political intervention and science in Latin America. Bulletin of the Atomic Scientists, 37(2):14–23, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [GS81].

Strom:1984:CIP

[Str84] Ted Strom. Commentary: International physicians met. Bulletin of the Atomic Scientists, 40(7):46, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stremmel:1986:LWP

[Str86] William J. Stremmel. Letter: World population. Bulletin of the Atomic Scientists, 42(7):65, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strong:1988:LOS

[Str88] Susan C. Strong. Letter: Open skies bill. Bulletin of the Atomic Scientists, 44(7):61–62, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lin88] and comment [Rob88].

Strasser:1990:LDT

[Str90] Marilyn Strasser. Letter: Don’t throw out that tinfoil. Bulletin of the Atomic Scientists, 46(3):55, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strekal:1993:RUW

[Str93] Oleg Strekal. Report: Ukraine: Westward ho! Bulletin of the Atomic Scientists, 49(9):10–11, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strekal:1994:NWR

[Str94] Oleg Strekal. No way to run an army. Bulletin of the Atomic Scientists, 50(1):34, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Strauss:2005:YC

[Str05a] Mark Strauss. A 60-year conspiracy. Bulletin of the Atomic Scientists, 61(6):4, November/December 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/6/4.

Strauss:2005:ED

[Str05b] Mark Strauss. Essence of a decision. Bulletin of the Atomic Scientists, 61(4):4, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/4/4.

Strauss:2005:NLE

[Str05c] Mark Strauss. A new look at Eisenhower. Bulletin of the Atomic Scientists, 61(4):69–70, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/4/69.

Strauss:2005:SF

[Str05d] Mark Strauss. Science friction. Bulletin of the Atomic Scientists, 61(5):4, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/5/4.

Strauss:2005:SAF

[Str05e] Mark Strauss. Some of all fears. Bulletin of the Atomic Scientists, 61(3):6, May/June 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/61/3/6.

Strauss:2005:WP

[Str05f] Mark Strauss. A wartime press. Bulletin of the Atomic Scientists, 61(5):70–71, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/5/70.

Strauss:2006:BCD

[Str06a] Mark Strauss. Behind closed doors. Bulletin of the Atomic Scientists, 62(3):4, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/3/4.

Strauss:2006:BM

[Str06b] Mark Strauss. Building materials. Bulletin of the Atomic Scientists, 62(2):4, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/2/4.

Strauss:2006:CNP

[Str06c] Mark Strauss. Crime and no punishment. Bulletin of the Atomic Scientists, 62(6):4, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/6/4.

Strauss:2006:FYL

[Str06d] Mark Strauss. Five years later. Bulletin of the Atomic Scientists, 62(5):4, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/5/4.

Strauss:2006:GB

[Str06e] Mark Strauss. Guns and butter. Bulletin of the Atomic Scientists, 62(1):65, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/1/65.1.

Strauss:2006:MT

[Str06f] Mark Strauss. A matter of timing. Bulletin of the Atomic Scientists, 62(4):4, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/4/4.

Strauss:2006:TA

[Str06g] Mark Strauss. Threat assessments. Bulletin of the Atomic Scientists, 62(1):4, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/62/1/4.

Strauss:2006:WIM

[Str06h] Mark Strauss. What’s in it for me? Bulletin of the Atomic Scientists, 62(4):63, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/4/63.

Strauss:2007:AA

[Str07a] Mark Strauss. Aging arsenals. Bulletin of the Atomic Scientists, 63(4):4, July 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/4/4.

Strauss:2007:EL

[Str07b] Mark Strauss. Enduring legacies. Bulletin of the Atomic Scientists, 63(3):4, May/June 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/3/4.

Strauss:2007:HP

[Str07c] Mark Strauss. Hungry for power. Bulletin of the Atomic Scientists, 63(5):4, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/63/5/4.

Strauss:2007:MS

[Str07d] Mark Strauss. Motivational seekers. Bulletin of the Atomic Scientists, 63(6):4, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/6/4.

Strauss:2007:TS

[Str07e] Mark Strauss. Taking stands. Bulletin of the Atomic Scientists, 63(2):4, March/April 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/2/4.

Strauss:2007:T

[Str07f] Mark Strauss. Time out? Bulletin of the Atomic Scientists, 63(1): 4, January/February 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/63/1/4.

Stanford:1980:CWW

[STRS80] G. S. Stanford, Laurence H. Tribe, David H. Remes, and John C. Stedman. Commentary: Will we perish if we publish? Bulletin of the Atomic Scientists, 36(6):62–64, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TR80, Sta80].

Szilard:1949:DSB

[STS^{+}49] Leo Szilard, The Franck Committee, Frederick Seitz, Hans Bethe, Harold C. Urey, Harrison Brown, Irving Langmuir, Vannevar Bush, and Leslie R. Groves. Did the Soviet bomb come sooner than expected? The atomic scientists: The wartime administrators. Bulletin of the Atomic Scientists, 5(10):262–264, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SGWAS:1947:TNS

[Stu47] Study Group of the Washington Association of Scientists. Toward a national science policy. Bulletin of the Atomic Scientists, 3(12): 357–358, 369, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Science. Group members are: G. Grobstein (Chair), J. M. Conly, I. Feister, L. B. Heilprin, H. Olken, F. J. Pratt, J. W. Rowen, I. Schocken, G. R. Silbiger, R. D. Steiehler, F. J. Weiss, and L. A. Wood.

SGWAS:1948:NSF

[Stu48] Study Group, Washington Association of Scientists. National Science Foundation: The Steelman report misses the point. Bulletin of the Atomic Scientists, 4(1):30–31, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stuhlinger:1969:APP

[Stu69a] Ernst Stuhlinger. Apollo: A pattern for problem solving. Bulletin of the Atomic Scientists, 25(7):79–83, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stuhlinger:1969:CCP

[Stu69b] Ernst Stuhlinger. Current comments: A prelude to space research. Bulletin of the Atomic Scientists, 25(3):24–27, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stunkel:1973:TS

[Stu73] Kenneth R. Stunkel. The technological solution. Bulletin of the Atomic Scientists, 29(7):42–44, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stubbs:1987:SSD

[Stu87] Eric Stubbs. Soviet strategic defense technology. Bulletin of the Atomic Scientists, 43(3):14–19, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stubbs:1989:BRB

[Stu89] Eric Stubbs. Book review: Soviet Ballistic Missile Defense and the Western Alliance, by David S. Yost. Bulletin of the Atomic Scientists, 45(6):46–47, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Subrahmanyam:1968:DPI

[Sub68] K. Subrahmanyam. Defense preparations in India and China. Bulletin of the Atomic Scientists, 24(5):28–33, May 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

SSRD:1970:RMA

[Sub70] Subcommittee on Science, Research, and Development. Reports: Mission agency support of basic research. Bulletin of the Atomic Scientists, 26(7):35–37, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Subrahmanyan:1977:NII

[Sub77] K. Subrahmanyan. The nuclear issue and international security. Bulletin of the Atomic Scientists, 33(2):17–21, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Subrahmanyan:1984:RCN

[Sub84] K. Subrahmanyan. Regional conflicts and nuclear fears. Bulletin of the Atomic Scientists, 40(5):16–19, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tho84c].

Subrahmanyan:1987:IVE

[Sub87] K. Subrahmanyan. International views: Eradicate the nuclear cult. Bulletin of the Atomic Scientists, 43(2):36–39, March 1987.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Subramanyam:1993:EON

[Sub93] K. Subramanyam. An equal-opportunity NPT. Bulletin of the Atomic Scientists, 49(5):37–39, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suess:1961:FRC

[Sue61] Hans E. Suess. Fuel residuals and climate. Bulletin of the Atomic Scientists, 17(9):374–375, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suess:1968:VHC

[Sue68] Hans E. Suess. Virus house: Comments and reminiscences. Bulletin of the Atomic Scientists, 24(6):36–39, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suedfeld:1982:CDC

[Sue82] Peter Suedfeld. Commentary: Doomsday clock. Bulletin of the Atomic Scientists, 38(3):46–47, March 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suica:1971:PDY

[Sui71] J. Salom Suica. Power development in Yugoslavia. Bulletin of the Atomic Scientists, 27(9):42–46, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sulzberger:1953:SS

[Sul53] Arthur Hays Sulzberger. “Smoke screen”. Bulletin of the Atomic Scientists, 9(2):47, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sullivan:1958:ISA

[Sul58a] Walter Sullivan. The IGY — scientific alliance in a divided world. Bulletin of the Atomic Scientists, 14(2):68–72, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sulzberger:1958:WSI

[Sul58b] C. L. Sulzberger. What others say: 2. implications of nuclear Christmas. Bulletin of the Atomic Scientists, 14(2):67, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Supek:1961:GEY

[Sup61] Ivan Supek. Guest editorial from Yugoslavia. Bulletin of the Atomic Scientists, 17(10):402, 431, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Supek:1967:NDE

[Sup67] Ivan Supek. New directions in education. Bulletin of the Atomic Scientists, 23(5):31–33, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sutton:1954:BBS

[Sut54] Constance Sutton. Books: The Study of Culture at a Distance, edited by Margaret Mead and Rhoda Metraux. Bulletin of the Atomic Scientists, 10(6):231–232, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sutton:1961:SV

[Sut61] George P. Sutton. Space vehicles. Bulletin of the Atomic Scientists, 17(5–6):194–199, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sutherland:1968:GSB

[Sut68a] Gordon Sutherland. Government and science in Britain and the U.S. Bulletin of the Atomic Scientists, 24(4):20–28, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sutton:1968:REB

[Sut68b] Samuel Sutton. Recent eddies in brain currents. Bulletin of the Atomic Scientists, 24(6):23–27, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suttmeier:1970:CCS

[Sut70] Richard P. Suttmeier. Communications: China’s scientists. Bulletin of the Atomic Scientists, 26(3):47–48, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sutherland:1977:LRM

[Sut77] Keith Sutherland. Letter: The real menace. Bulletin of the Atomic Scientists, 33(4):2, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Flo76a, Pen77b].

Sutherland:1983:CHF

[Sut83] Keith Sutherland. Commentary: The Harris findings. Bulletin of the Atomic Scientists, 39(1):52, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kal82h].

Suttmeier:1984:STC

[Sut84] Richard P. Suttmeier. Science and technology in China: New conflicts in the research environment. Bulletin of the Atomic Scientists, 40(8):7s–11s [various between 24–25], October 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sutherland:1989:LDB

[Sut89] Norman E. Sutherland. Letter: Don’t bring the boys home. Bulletin of the Atomic Scientists, 45(9):50–51, November 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RL89].

Sutcliffe:1991:LDW

[Sut91] William G. Sutcliffe. Letter: Dismantling warheads serious business. Bulletin of the Atomic Scientists, 47(1):46, January 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [NA90a].

Suter:1995:RNT

[Sut95] Keith Suter. Report: Nuclear testing: Paradise lost. Bulletin of the Atomic Scientists, 51(5):13–14, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suter:1996:HP

[Sut96a] Keith Suter. Homefront politics. Bulletin of the Atomic Scientists, 52(6):42, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suter:1996:RTR

[Sut96b] Keith Suter. Report: Treaty of Rarotonga: U.S. signs on at last. Bulletin of the Atomic Scientists, 52(2):12–13, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suter:1997:NHE

[Sut97] Keith Suter. Nobel hears East Timor. Bulletin of the Atomic Scientists, 53(1):47–48, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suter:2001:IIW

[Sut01] Keith Suter. Indonesia: Independence for West Papua? Bulletin of the Atomic Scientists, 57(3):16–18, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/16.

Suter:2004:ASS

[Sut04] Keith Suter. Australia: Sheriff of the South Pacific. Bulletin of the Atomic Scientists, 60(1):18–19, January 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/1/18.

Sutyagin:2016:RU

[Sut16] Igor Sutyagin. Russia’s underwater “doomsday drone”: Science fiction, but real danger. Bulletin of the Atomic Scientists, 72(4): 243–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suzuki:1984:PAT

[Suz84] Sunao Suzuki. Public attitudes toward peace. Bulletin of the Atomic Scientists, 40(2):27–31, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Suzuki:2011:DZR

[Suz11] Tatsujiro Suzuki. Deconstructing the zero-risk mindset: The lessons and future responsibilities for a post-Fukushima nuclear Japan. Bulletin of the Atomic Scientists, 67(5):9–18, September/ October 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/67/
5/9.

Sandman:1986:SSS

[SV86] Peter M. Sandman and JoAnn M. Valenti. Scared stiff — or scared into action. Bulletin of the Atomic Scientists, 42(1):12– 16, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [AC86, Cal86a].

Schmidt:1977:LDW

[SvH77] Fred H. Schmidt and Frank von Hippel. Letter: Deception by whom? Bulletin of the Atomic Scientists, 33(5):4, May 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH77b].

Seitz:1956:GCD

[SW56] Frederick Seitz and Eugene Wigner. On the Geneva Conference: a dissenting opinion. Bulletin of the Atomic Scientists, 12(1):23– 24, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also [Lap56c].

Seitz:1959:GPA

[SW59] Frederick Seitz and Eugene P. Wigner. Geneva, 1958. pure and applied nuclear physics in East and West. Bulletin of the Atomic Scientists, 15(3):127–131, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1980:POR

[SW80a] Alice Kimball Smith and Charles Weiner. Part one: Robert Oppenheimer: Letters and recollections. Bulletin of the Atomic Scientists, 36(5):19–27, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Smith:1980:PTR

[SW80b] Alice Kimball Smith and Charles Weiner. Part two: Robert Oppenheimer: the Los Alamos years. Bulletin of the Atomic Scientists, 36(6):11–17, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Salama:2007:ANE

[SW07] Sammy Salama and Heidi Weber. Arab nuclear envy. Bulletin of the Atomic Scientists, 63(5):??, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/5/44.

Selgelid:2010:RSP

[SW10] Michael J. Selgelid and Lorna Weir. Reflections on the synthetic production of poliovirus. Bulletin of the Atomic Scientists, 66(3):

1–9, May 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/66/
3/1.

Stephens:2016:CCT

[SW16] Jennie C. Stephens and Elizabeth J. Wilson. Climate change, technological innovation: a US response. Bulletin of the Atomic Scientists, 72(1):4–6, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Simon:2021:AUS

[SW21] Maxwell Simon and Sam Wilson. Avoiding an unintentional space war: Lessons from Cold War nuclear diplomacy. Bulletin of the Atomic Scientists, 77(3):140–145, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1912243.

Swartout:1949:ASC

[Swa49] John A. Swartout. AEC security clearance: The scientists’ viewpoint. Bulletin of the Atomic Scientists, 5(12):337–339, 351, December 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Swanson:1966:ICA

[Swa66] Don R. Swanson. On improving communication among scientists. Bulletin of the Atomic Scientists, 22(2):8–12, February 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Swain:1986:LHS

[Swa86] David L. Swain. Letter: Hiroshima survivors’ stories. Bulletin of the Atomic Scientists, 42(3):55, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [She85e].

Simmons:1984:CTA

[SWC^{+}84] Adele Simmons, Dennis Wallick, Philip Carman, Joseph S. Nye, Jr., and Henry T. Nash. Commentary: On ‘Thinking about thinking about the unthinkable’. Bulletin of the Atomic Scientists, 40 (2):51–56, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Nas83].

Sweet:1977:ONP

[Swe77] William Sweet. The opposition to nuclear power in Europe. Bulletin of the Atomic Scientists, 33(10):40–47, December 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Dua78, Str78].

Sweet:1978:CWS

[Swe78a] William Sweet. Commentary: From Washington: Seabrook, the “Clams” and the Commission. Bulletin of the Atomic Scientists, 34(8):53–54, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1978:UIS

[Swe78b] William Sweet. The U.S.–India safeguards dispute. Bulletin of the Atomic Scientists, 34(6):50–52, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1979:CWU

[Swe79a] William Sweet. Commentary from Washington: Unresolved: the front end of nuclear waste disposal. Bulletin of the Atomic Scientists, 35(5):44–48, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1979:WBN

[Swe79b] William Sweet. From Washington: Building up to negotiate down? Bulletin of the Atomic Scientists, 35(8):9–10, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1980:CWR

[Swe80] William Sweet. Commentary from Washington: Rating Carter on proliferation. Bulletin of the Atomic Scientists, 36(1):56–57, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1981:CMR

[Swe81] William Sweet. Commentary: On ‘The mouse that roared’. Bulletin of the Atomic Scientists, 37(6):62–63, June/July 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Shu81, Pai81a].

Sweet:1983:E

[Swe83] William Sweet. The 1982 election. Bulletin of the Atomic Scientists, 39(1):56–57, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1985:ALB

[Swe85] William Sweet. Another look at Deadly Gambits. Bulletin of the Atomic Scientists, 41(5):49–50, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1986:BRB

[Swe86] William Sweet. Book reviews: The Green Factor in German Politics, by Gerd Langguth; European Peace Movements and the Future of the Western Alliance, edited by Walter Laqueur; Shattering Europe’s Defense Consensus: The Antinuclear Protest Movement and the Future of NATO, edited by James E. Dougherty and Robert L. Pfaltzgraff, Jr. Bulletin of the Atomic Scientists, 42(6):52–54, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1987:LRS

[Swe87a] William Sweet. Letter: Radar and Star Wars. Bulletin of the Atomic Scientists, 43(3):52, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi87c].

Sweet:1987:PSS

[Swe87b] William Sweet. Perspective: Scientists shoot down Star Wars. Bulletin of the Atomic Scientists, 43(6):7–9, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1988:BRB

[Swe88] William Sweet. Book review: Hope in Hard Times: America’s Peace Movement in the Reagan Era, by Paul Rogat Loeb. Bulletin of the Atomic Scientists, 44(2):50–51, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Sweet:1989:BRB

[Swe89a] William Sweet. Book review: From Protest to Policy: Beyond the Freeze to Common Security, by Pam Solo. Bulletin of the Atomic Scientists, 45(2):43–44, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bro89a].

Sweet:1989:LDG

[Swe89b] William Sweet. Letter: Data gone. Bulletin of the Atomic Scientists, 45(3):44, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kal89].

Sweet:1993:BRU

[Swe93] William Sweet. Book reviews: Uncertainties: Heisenberg’s War: The Secret History of the German Bomb, by Thomas Powers; Uncertainty: The Life and Science of Werner Heisenberg, by David C. Cassidy. Bulletin of the Atomic Scientists, 49(7):50– 52, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wei93].

Sweet:2002:BLN

[Swe02] William Sweet. The Bohr letters: No more uncertainty. Bulletin of the Atomic Scientists, 58(3):20–27, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/3/20.

Swing:1983:LS

[Swi83] John Temple Swing. Law of the Sea. Bulletin of the Atomic Scientists, 39(5):14–19, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Swonger:1986:LCT

[Swo86] James W. Swonger. Letter: Concessionist tracts. Bulletin of the Atomic Scientists, 42(4):53, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Symington:1950:ICD

[Sym50] W. Stuart Symington. The importance of civil defense planning. Bulletin of the Atomic Scientists, 6(8–9):231–233, August/ September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Strauss:1954:AMD

[SZC54] Lewis L. Strauss, Eugene M. Zuckert, and Joseph Campbell. AEC majority decision: 29 June 1954. Bulletin of the Atomic Scientists, 10(7):275–277, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1947:ABP

[Szi47a] Leo Szilard. Atomic bombs and the postwar position of the United States in the world — 1945. Bulletin of the Atomic Scientists, 3 (12):351–353, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1947:CC

[Szi47b] Leo Szilard. Calling for a crusade. Bulletin of the Atomic Scientists, 3(4–5):102–106, 125, April/May 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1947:CED

[Szi47c] Leo Szilard. Comment to the Editors by Dr. Szilard. Bulletin of the Atomic Scientists, 3(12):350, 353, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Szi47d, Rab47o].

Szilard:1947:LS

[Szi47d] Leo Szilard. Letter to Stalin. Bulletin of the Atomic Scientists, 3(12):348–349, 376, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Rab47o, Szi47c].

Szilard:1949:AFS

[Szi49a] Leo Szilard. The AEC Fellowships: Shall we yield or fight? Bulletin of the Atomic Scientists, 5(6–7):177–178, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Szi63a].

Szilard:1949:SWF

[Szi49b] Leo Szilard. Shall we face the facts? Bulletin of the Atomic Scientists, 5(10):269–273, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1950:CWI

[Szi50a] Leo Szilard. Can we have international control of atomic energy? Bulletin of the Atomic Scientists, 6(1):9–12, 16, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1950:DDD

[Szi50b] Leo Szilard. The diary of Dr. Davis. Bulletin of the Atomic Scientists, 6(2):51–57, February 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1954:SR

[Szi54] Leo Szilard. Security risk. Bulletin of the Atomic Scientists, 10(10):384–386, 398, December 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1955:DPP

[Szi55a] Leo Szilard. Disarmament and the problem of peace. Bulletin of the Atomic Scientists, 11(8):297–307, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on Secrecy, Security, and Loyalty. See comment [Fow55].

Szilard:1955:FSP

[Szi55b] Leo Szilard. The first step to peace. Bulletin of the Atomic Scientists, 11(3):104, March 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1960:HLB

[Szi60a] Leo Szilard. How to live with the bomb and survive — the possibility of a Pax Russo–Americana in the long-range rocket stage of the so-called atomic stalemate. Bulletin of the Atomic Scientists, 16(2):59–73, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Szi63b].

Szilard:1960:LEB

[Szi60b] Leo Szilard. Letter to the Editor: The Berlin crisis. Bulletin of the Atomic Scientists, 16(5):144, 193, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1960:SS

[Szi60c] Leo Szilard. To stop or not to stop. Bulletin of the Atomic Scientists, 16(3):82–84, 108, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Szi63c].

Szilard:1961:LEA

[Szi61a] Leo Szilard. Letter to the Editor: American Society of Newspaper Editors and President Kennedy. Bulletin of the Atomic Scientists, 17(5–6):168, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1961:MC

[Szi61b] Leo Szilard. The mined cities. Bulletin of the Atomic Scientists, 17(10):407–412, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1962:WRW

[Szi62a] Leo Szilard. Are we on the road to war? Bulletin of the Atomic Scientists, 18(4):23–30, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1962:CD

[Szi62b] Leo Szilard. The Council’s dilemma. Bulletin of the Atomic Scientists, 18(10):50, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1963:AFS

[Szi63a] Leo Szilard. The AEC Fellowships: Shall we yield or fight? In Grodzins and Rabinowitch [GR63], pages 410–413. LCCN D842 .B78. Reprint of [Szi49a].

Szilard:1963:HLB

[Szi63b] Leo Szilard. How to live with the bomb — and survive. In Grodzins and Rabinowitch [GR63], pages 217–243. LCCN D842 .B78. Reprint of [Szi60a].

Szilard:1963:TT

[Szi63c] Leo Szilard. To test or not to test. In Grodzins and Rabinowitch [GR63], pages 342–347. LCCN D842 .B78. Reprint of [Szi60c].

Szilard:1964:MDV

[Szi64] Leo Szilard. “Minimal deterrent” vs. saturation parity. Bulletin of the Atomic Scientists, 20(3):6–12, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1965:SBS

[Szi65] Leo Szilard. The “Sting of the bee” in saturation parity. Bulletin of the Atomic Scientists, 21(3):8–13, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Szilard:1971:CLC

[Szi71] Gertrud Weiss Szilard. Communications: Lunar craters. Bulletin of the Atomic Scientists, 27(4):4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thompson:1994:LSA

[TA94] Peter Thompson and David Albright. Letters: South Africa’s nuclear motive. Bulletin of the Atomic Scientists, 50(6):57, November/December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taeuber:1951:CTP

[Tae51] Irene B. Taeuber. Culture, technology, and population change. Bulletin of the Atomic Scientists, 7(7–8):206–208, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Takahashi:1984:RH

[Tak84] Shinji Takahashi. Relief for the Hibakusha. Bulletin of the Atomic Scientists, 40(8):25–26, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Takubo:2011:NCU

[Tak11] Masa Takubo. Nuclear or not? The complex and uncertain politics of Japan’s post-Fukushima energy policy. Bulletin of the Atomic Scientists, 67(5):19–26, September/October 2011. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/67/5/19.

Takubo:2017:CJM

[Tak17] Masa Takubo. Closing Japan’s Monju fast breeder reactor: The possible implications. Bulletin of the Atomic Scientists, 73(3): 182–187, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1315040.

Talensky:1965:ASD

[Tal65] N. Talensky. Antimissile systems and disarmament. Bulletin of the Atomic Scientists, 21(2):26–29, February 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [All65].

Talbot:2018:EIP

[Tal18] Brent J. Talbot. Eliminating ICBMs — as part of a 21stcentury deterrence strategy. Bulletin of the Atomic Scientists, 74(1):52–59, 2018. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1413067.

Tamplin:1969:CFI

[Tam69] Arthur R. Tamplin. Comments: Fetal and infant mortality. Bulletin of the Atomic Scientists, 25(10):22–29, December 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e] and reply [Ste69d].

Tamers:1971:LGM

[Tam71a] Murry A. Tamers. The language gap: Must we all speak English? Bulletin of the Atomic Scientists, 27(3):38–40, March 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bru72].

Tammen:1971:LWN

[Tam71b] Ronald L. Tammen. Letter from Washington: New strategy, new threats: The fiscal 1972 military budget. Bulletin of the Atomic Scientists, 27(5):42–44, May 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tamplin:1971:IRC

[Tam71c] Arthur R. Tamplin. Issues in the radiation controversy. Bulletin of the Atomic Scientists, 27(7):25–27, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tamplin:1973:BRF

[Tam73a] Arthur R. Tamplin. The BEIR report: A focus on issues. Bulletin of the Atomic Scientists, 29(5):19–20, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kat73] and response [Tam73b].

Tamplin:1973:CR

[Tam73b] Arthur R. Tamplin. Communications: Response. Bulletin of the Atomic Scientists, 29(10):3, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tam73a, Kat73].

Tamplin:1974:RVP

[Tam74] Arthur R. Tamplin. Random violence in the Pacific: Estimate of dosage to man of cesium-136 in weapons test fallout. Bulletin of the Atomic Scientists, 30(3):10–14, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tanner:1983:KDA

[Tan83] Daniel Tanner. Knowledge divided against itself. Bulletin of the Atomic Scientists, 39(3):34–38, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tanaka:1984:EMN

[Tan84] Naoki Tanaka. The economy: military and non-military aspects. Bulletin of the Atomic Scientists, 40(2):24–27, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tanaka:1988:PGS

[Tan88] Yuki Tanaka. Poison gas: the story Japan would like to forget. Bulletin of the Atomic Scientists, 44(8):10–19, October 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tannenwald:2006:TS

[Tan06] Nina Tannenwald. A taboo subject. Bulletin of the Atomic Scientists, 62(3):64, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/3/64.

Tarasenko:1993:RMS

[Tar93] Maxim Tarasenko. Report: Military space: Twinkle, twinkle little Topaz. Bulletin of the Atomic Scientists, 49(6):11–13, July/ August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tarapore:2023:CRW

[Tar23] Arzan Tarapore. Conditional restraint: Why the India– Pakistan Kargil War is not a case of nuclear deterrence. Bulletin of the Atomic Scientists, 79(6):388–392, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2266943.

Taschdjian:1951:PFP

[Tas51] Edgar Taschdjian. Problems of food production. Bulletin of the Atomic Scientists, 7(7–8):209–213, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tatarian:1978:MNP

[Tat78] Roger Tatarian. A multinational news pool. Bulletin of the Atomic Scientists, 34(7):54–56, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taunton:1985:LTR

[Tau85] Joel Taunton. Letter: Tax resistance. Bulletin of the Atomic Scientists, 41(11):64, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1947:TSH

[Tay47] Glen H. Taylor. Is there some hope for world government? Bulletin of the Atomic Scientists, 3(10):289, 304, October 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1952:LEA

[Tay52] Hugh S. Taylor. Letter to the Editor: Attitude of the Catholic Church toward science. Bulletin of the Atomic Scientists, 8(6): 209, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1956:LEP

[Tay56] Wm. Palmer Taylor. Letter to the Editor: Proposals for control of nuclear weapons. Bulletin of the Atomic Scientists, 12(10): 381, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1958:IA

[Tay58] Harold Taylor. The intellectual in action. Bulletin of the Atomic Scientists, 14(9):368–373, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1962:LEC

[Tay62] Wm. Palmer Taylor. Letter to the Editor: Clash of symbols. Bulletin of the Atomic Scientists, 18(3):39, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1969:LED

[Tay69] William Palmer Taylor. Letter to the Editor: Dyson views challenged. Bulletin of the Atomic Scientists, 25(9):46, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys69a].

Taylor:1974:CPN

[Tay74] Byron M. Taylor. Communications: President Nixon’s plan. Bulletin of the Atomic Scientists, 30(2):3, 45, February 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1976:LMI

[Tay76] William Palmer Taylor. Letter: Military inefficiency. Bulletin of the Atomic Scientists, 32(7):2, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1981:CSE

[Tay81] Theodore B. Taylor. Commentary: Solar energy and peace. Bulletin of the Atomic Scientists, 37(2):6–7, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1986:EGN

[Tay86] Theodore B. Taylor. Endless generations of nuclear weapons. Bulletin of the Atomic Scientists, 42(9):12–15, November 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sha87b].

Taylor:1987:BRB

[Tay87] Theodore B. Taylor. Book review: Making Europe Unconquerable: The Potential of Civilian-Based Deterrence and Defense, by Gene Sharp. Bulletin of the Atomic Scientists, 43(1):54–56, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tsc87].

Taylor:1989:GCT

[Tay89] Theodore B. Taylor. Go cold turkey. Bulletin of the Atomic Scientists, 45(6):26–27, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1991:BRD

[Tay91] Theodore B. Taylor. Book review: Dumping on Nevada: Site Unseen, by Gerald Jacob. Bulletin of the Atomic Scientists, 47 (6):41–42, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1992:JUE

[Tay92] Theodore B. Taylor. Just unplug ’em. Bulletin of the Atomic Scientists, 48(4):27–28, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:1996:GOC

[Tay96] Theodore B. Taylor. Guest opinion: Circles of destruction. Bulletin of the Atomic Scientists, 52(1):4, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-17773681.
html.

Taylor:1998:BRS

[Tay98] Bryan Taylor. Book review: Symbolic Los Alamos: Secret Mesa: Inside Los Alamos National Laboratory, by Jo Ann Shroyer. Bulletin of the Atomic Scientists, 54(3):65–67, May/June 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taylor:2021:LDS

[Tay21] Terence Taylor. Lessons to be drawn from the search for Iraqi WMD. Bulletin of the Atomic Scientists, 77(4):180–183, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1941553.

Thompson:1970:RGT

[TB70] Theos J. Thompson and William R. Bibb. Response to Gofman and Tamplin: The AEC position. Bulletin of the Atomic Scientists, 26(7):9–12, 48, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TG70, Pau70].

Tilden:2023:NNN

[TB23] Jay A. Tilden and Dallas Boyd. Nerds, ninjas, and neutrons: The story of the Nuclear Emergency Support Team. Bulletin of the Atomic Scientists, 79(2):87–94, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2178179.

Taylor:1985:CNW

[TBBS85] John Taylor, Ralph Balent, John Barton, and Thomas Shea. Converting nuclear weapons to peaceful use. Bulletin of the Atomic Scientists, 41(2):35–38, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Trefall:1983:CNF

[TC83] Harold Trefall and Michael Carver. Commentary: No first use: a view from Europe. Bulletin of the Atomic Scientists, 39(6): 53, 55–56, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Car83c].

Taylor:1989:DNF

[TD89] Theodore B. Taylor and Sidney Drell. [Debate: a nuclear-free millennium]. Bulletin of the Atomic Scientists, 45(6):29–31, July/ August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tebicke:1985:ASA

[Teb85] Haile Lul Tebicke. Africa: Sustainable agriculture: an Ethiopian view. Bulletin of the Atomic Scientists, 41(8):39–42, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teeple:1965:RAC

[Tee65] John B. Teeple. Recent arms control research in Europe. Bulletin of the Atomic Scientists, 21(8):37–39, October 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1946:SWP

[Tel46a] Edward Teller. Scientists in war and peace. Bulletin of the Atomic Scientists of Chicago, 1(6):10–11, March 1, 1946. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1946:SDR

[Tel46b] Edward Teller. The State Dep’t report — ‘a ray of hope’. Bulletin of the Atomic Scientists, 1(8):10, 13, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1946:SAA

[Tel46c] Edward Teller. A suggested amendment to the Acheson Report. Bulletin of the Atomic Scientists, 1(12):5, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1947:AST

[Tel47a] Edward Teller. The atomic scientists have two responsibilities. Bulletin of the Atomic Scientists, 3(12):355–356, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Tel63].

Teller:1947:HEM

[Tel47b] Edward Teller. The high-energy machines: their design and application. Bulletin of the Atomic Scientists, 3(7):171–172, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1947:HDA

[Tel47c] Edward Teller. How dangerous are atomic weapons? Bulletin of the Atomic Scientists, 3(2):35–36, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1947:PDF

[Tel47d] Edward Teller. Politics, domestic and foreign. Bulletin of the Atomic Scientists, 3(3):100, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1948:CDW

[Tel48a] Edward Teller. Comments on the “Draft of a World Constitution”. Bulletin of the Atomic Scientists, 4(7):204, July 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com48a].

Teller:1948:FYA

[Tel48b] Edward Teller. The first year of the Atomic Energy Commission. Bulletin of the Atomic Scientists, 4(1):5–6, January 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1949:BRL

[Tel49] Edward Teller. Book review: Lang’s Early Tales of the Atomic Age, by Daniel Lang, New York: Doubleday, 1948. Bulletin of the Atomic Scientists, 5(2):64, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1950:BL

[Tel50] Edward Teller. Back to the laboratories. Bulletin of the Atomic Scientists, 6(3):71–72, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Telkes:1951:FUS

[Tel51a] Maria Telkes. Future uses of solar energy. Bulletin of the Atomic Scientists, 7(7–8):217–219, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ma’52].

Teller:1951:S

[Tel51b] Edward Teller. To Sagittarius. Bulletin of the Atomic Scientists, 7 (1):22, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1954:LAA

[Tel54] Edward Teller. Letter: The accomplishments of Los Alamos. Bulletin of the Atomic Scientists, 10(9):359, November 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1956:CTR

[Tel56] Edward Teller. Controlled thermonuclear reactions. Bulletin of the Atomic Scientists, 12(7):271–272, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1957:JND

[Tel57a] Edward Teller. John von Neumann [28 December 1903–8 February 1957]. Bulletin of the Atomic Scientists, 13(4):150–151, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1957:NNWb

[Tel57b] Edward Teller. The nature of nuclear warfare. Bulletin of the Atomic Scientists, 13(5):162–165, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Air Force Magazine, January 1957. See comment [Fer57c].

Teller:1958:RC

[Tel58] Edward Teller. The Russian challenge. Bulletin of the Atomic Scientists, 14(2):87–89, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1960:IP

[Tel60] Edward Teller. The issue of peace. Bulletin of the Atomic Scientists, 16(6):201–204, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1963:TRS

[Tel63] Edward Teller. The two responsibilities of scientists. In Grodzins and Rabinowitch [GR63], pages 121–124. LCCN D842 .B78. Reprint of [Tel47a].

Teller:1966:RAS

[Tel66a] Edward Teller. The role of applied science. Bulletin of the Atomic Scientists, 22(3):15–19, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tellez:1966:RCA

[Tel66b] Theresa Tellez. Reports: The crisis of Argentine science. Bulletin of the Atomic Scientists, 22(10):32–34, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tellez:1968:MSN

[Tél68] Theresa Téllez. Mexican science: A new era? Bulletin of the Atomic Scientists, 24(4):46–48, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1971:EBS

[Tel71] Edward Teller. The era of big science. Bulletin of the Atomic Scientists, 27(4):34–36, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tellez:1973:TSS

[Tel73] Theresa Tellez. Technical solutions: Science, technology and the matter of choice. Bulletin of the Atomic Scientists, 29(8):55–57, October 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tellefsen:1975:LMN

[Tel75] Olaf Tellefsen. Letters: Man and nature. Bulletin of the Atomic Scientists, 31(2):3–4, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gol75].

Tellez:1976:BST

[Tel76] Theresa Tellez. Books: Science in the third world: Science Development: The Building of Science in Less Developed Countries, by Michael J. Moravcsik. Bulletin of the Atomic Scientists, 32(3): 63–64, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1982:DMA

[Tel82] Edward Teller. Dangerous myths about nuclear arms. Reader’s Digest, ??(??):139–143, November 1982. CODEN RDIGAU. ISSN 0034-0375. See comment [vH83a], in which this article is reprinted.

Teller:1991:SNG

[Tel91] Edward Teller. START: Nuclear glasnost. Bulletin of the Atomic Scientists, 47(9):34–35, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sho92a].

Tenney:1969:LEW

[Ten69] Fred H. Tenney. Letter to the Editor: What does this mean? Bulletin of the Atomic Scientists, 25(9):45–46, November 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys69a].

Teng:1975:BBI

[Ten75] Lee C. Teng. Books: Creativity and Intuition: A Physicist Looks East and West, by Hideki Yukawa, translated by John Bester. Bulletin of the Atomic Scientists, 31(3):48–49, March 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Terrill:1958:RMP

[Ter58] James G. Terrill, Jr. Radiation and man: Peacetime radiation: Some public health aspects of radioactive wastes. Bulletin of the Atomic Scientists, 14(1):44–45, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Terral:1964:KR

[Ter64] Rufus Terral. To kill a river. Bulletin of the Atomic Scientists, 20 (7):35–37, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tertrais:2004:NPF

[Ter04] Bruno Tertrais. Nuclear policy: France stands alone. Bulletin of the Atomic Scientists, 60(4):48–55, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/48.

Toornstra:1981:CNB

[TFM81] Dick Toornstra, Berge Furre, and Alva Myrdal. Commentary: Neutron bomb: reactions from Europe. Bulletin of the Atomic Scientists, 37(9):60–62, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tamplin:1970:TGP

[TG70] Arthur W. Tamplin and John W. Gofman. Tamplin–Gofman, Pauling and the AEC: The radiation effects controversy. Bulletin of the Atomic Scientists, 26(7):2, 5–8, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [TB70, Tot70, Rum71b, Cow71] and [Pau70].

Thant:1963:ELO

[Tha63] U. Thant. [Extermination of life on our small planet?]. Bulletin of the Atomic Scientists, 19(7):23, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thackray:1978:LDL

[Tha78] M. Thackray. Letter: DNA and the law. Bulletin of the Atomic Scientists, 34(2):6–7, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thakur:2000:WUS

[Tha00] Ramesh Thakur. Will the United States take the lead? Bulletin of the Atomic Scientists, 56(1):6, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/1/6.

Editors:1948:MJC

[The48] The Editors. In memoriam: Joyce Clennam Stearns. Bulletin of the Atomic Scientists, 4(8):235, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

TEMB:1979:P

[The79] The Earl Mountbatten of Burma. Perspective. Bulletin of the Atomic Scientists, 35(8):1, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Theorin:1984:PDM

[The84] Maj Britt Theorin. Perspective: Deploy missiles, delay peace. Bulletin of the Atomic Scientists, 40(4):2–3, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

TE:1988:ENL

[The88] The Editors. [Editorial note: Latest scram at Savannah reactors]. Bulletin of the Atomic Scientists, 44(10):7, December 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

TE:1989:ENE

[The89a] The Editors. Editor’s note: Exporting trouble — West Germany’s freewheeling nuclear business. Bulletin of the Atomic Scientists, 45(4):55, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cha89b].

TE:1989:EC

[The89b] The Editors. From the Editors: Collaborators. Bulletin of the Atomic Scientists, 45(3):2, April 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

TE:1989:ELN

[The89c] The Editors. From the Editors: Those lying numbers. Bulletin of the Atomic Scientists, 45(2):2, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

TE:1989:ETB

[The89d] The Editors. From the Editors: Today’s Bulletin. Bulletin of the Atomic Scientists, 45(8):3, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

TE:1989:OBR

[The89e] The Editors. Omission: [book review: Edgar Snow, by John M. Hamilton]. Bulletin of the Atomic Scientists, 45(6):54, July/ August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pow89b].

TE:1989:RSL

[The89f] The Editors. Reagan’s security legacy. Bulletin of the Atomic Scientists, 45(1):5, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [McK89].

TE:1990:ENP

[The90] The Editors. [Editorial note: Peter (Pyotr) Leonidovich Kapitsa, Soviet scientist]. Bulletin of the Atomic Scientists, 46(3):26, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thirring:1946:GAE

[Thi46] Hans Thirring. Die Geschichte der Atombombe; mit einer elementaren Einführung in die Atomphysik auf Grund der Originalliteratur gemeinverständlich dargestellt. (German) [The history of the atomic bomb, with an elementary introduction to atomic physics according to the original literature, understandably presented]. Wissenschaft für jedermann; Phönix-Bücherel. “Neues Österreich” Zeitungs- und Verlagsgesellschaft, Wien, Austria, 1946. 149 pp. LCCN QC778 .T47.

Thimann:1947:RBW

[Thi47] Kenneth V. Thimann. The role of biologists in warfare. Bulletin of the Atomic Scientists, 3(8):211–212, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thirring:1950:SB

[Thi50] Hans Thirring. The super bomb. Bulletin of the Atomic Scientists, 6(3):69–70, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). English translation of a chapter of the author’s book [Thi46] (reviewed in the Bulletin in April 1948). According to [Har10, page 203] Its English translation was delayed until after President Truman’s order of 1 February 1950 directing the Atomic Energy Commission (AEC) to work on the super bomb.

Thirring:1952:IWU

[Thi52] Hans Thirring. Is it wise to use uranium for power? Bulletin of the Atomic Scientists, 8(6):171, 205–206, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thirring:1953:WA

[Thi53] Hans Thirring. Who is an aggressor? Bulletin of the Atomic Scientists, 9(3):68–72, 96, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thirring:1955:CWS

[Thi55] Hans Thirring. Can we sleep soundly in the shadow of the hydrogen bomb? Bulletin of the Atomic Scientists, 11(2):59–61, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thirring:1958:CPW

[Thi58] Hans Thirring. Comments on Pugwash: From the West: Perils from war vs. those from tests. Bulletin of the Atomic Scientists, 14(3):121–122, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thirring:1959:EAS

[Thi59] Hans Thirring. Education for the age of science. Bulletin of the Atomic Scientists, 15(7):292–297, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomson:1953:AUC

[Tho53] Sir George Thomson. Anglo–U.S. cooperation on atomic energy. Bulletin of the Atomic Scientists, 9(2):46–48, March 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomson:1957:NIR

[Tho57] Sir George Thomson. The new industrial resolution. Bulletin of the Atomic Scientists, 13(1):9–12, January 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thorstenson:1962:BBC

[Tho62] Joel S. Thorstenson. Books: War and the Christian Conscience, by Paul Ramsey. Bulletin of the Atomic Scientists, 18(4):41–42, April 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomas:1975:LNA

[Tho75] William A. Thomas. Letter: Neglected arenas. Bulletin of the Atomic Scientists, 31(10):4, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day75o].

Thompson:1977:LLC

[Tho77] Thurston A. Thompson. Letter: Let’s catch up. Bulletin of the Atomic Scientists, 33(8):2, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gou78a, La 78, Fel77g].

Thomas:1979:FRB

[Tho79a] Stephen M. Thomas. Film review: Survival … or Suicide. Bulletin of the Atomic Scientists, 35(7):48–49, September 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomasson:1979:DLD

[Tho79b] W. A. Thomasson. Deadly legacy: dioxin and the Vietnam veteran. Bulletin of the Atomic Scientists, 35(5):15–19, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DD79].

Thomasson:1979:RDR

[Tho79c] W. A. Thomasson. Recombinant DNA and regulating uncertainty. Bulletin of the Atomic Scientists, 35(10):26–32, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomas:1980:CNE

[Tho80a] Urs P. Thomas. Commentary: Is nuclear energy necessary? Bulletin of the Atomic Scientists, 36(10):58, December 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei80a].

Thomasson:1980:BRBc

[Tho80b] W. A. Thomasson. Book review: Report to the President, by the Toxic Substances Strategy Committee. Bulletin of the Atomic Scientists, 36(7):42, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomasson:1980:BRBa

[Tho80c] W. A. Thomasson. Book review: The Pendulum and the Toxic Cloud. The Course of Dioxin Contamination, by Thomas Whiteside. Bulletin of the Atomic Scientists, 36(1):54, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomasson:1980:BRBb

[Tho80d] W. A. Thomasson. Book reviews: Colonies in Space, by T. A. Heppenheimer; Toward Distant Suns, by T. A. Heppenheimer; Worlds Beyond: The Everlasting Frontier, edited by Larry Geis and Fabrice Florin; Living in Outer Space, by George S. Robinson; Star Trek: The Endless Migration, by Jerome Clayton Glenn and George S. Robinson. Bulletin of the Atomic Scientists, 36(6): 53–55, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomasson:1980:CP

[Tho80e] W. A. Thomasson. The Citizens Party 1980. Bulletin of the Atomic Scientists, 36(7):25–28, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomasson:1980:CCL

[Tho80f] W. A. Thomasson. Commentary: The candidates 1980 …. Bulletin of the Atomic Scientists, 36(8):13–19, October 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomasson:1980:RTB

[Tho80g] W. A. Thomasson. The reading table: From Genesis to Genocide. The Meaning of Human Nature and the Power of Behavior Control, by Stephen L. Chorover; On Human Nature, by Edward O. Wilson; A Nation of Guinea Pigs. The Unknown Risks of Chemical Technology, by Marshall S. Shapo; The Biology of Peace and War. Men, Animals, and Aggression, by Irenaus Eibl-Eibesfeldt, translated by Eric Mosbacher; Cloning. A Biologist Reports, by Robert Gilmore McKinnell; The Recombinant DNA Debate, edited by David A. Jackson and Stephen P. Stich; A Double Image of the Double Helix. The Recombinant DNA Debate, by Clifford Grobstein. Bulletin of the Atomic Scientists, 36(2):50– 53, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thompson:1980:CIR

[Tho80h] Gordon Thompson. Commentary: The Iraqi reactor controversy. Bulletin of the Atomic Scientists, 36(9):57, November 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thorsson:1980:PDV

[Tho80i] Inga Thorsson. Perspective: Disavowing violence. Bulletin of the Atomic Scientists, 36(7):1, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomas:1981:CNC

[Tho81a] Raju G. C. Thomas. Commentary: The nuclear club and affirmative action. Bulletin of the Atomic Scientists, 37(5):60–62, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thompson:1981:CLW

[Tho81b] Aileen P. Thompson. Commentary: A livable world. Bulletin of the Atomic Scientists, 37(7):64, August/September 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel81e].

Thompson:1981:EL

[Tho81c] E. P. Thompson. The END of the line. Bulletin of the Atomic Scientists, 37(1):6–13, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thorsson:1982:SDD

[Tho82] Inga Thorsson. Study on disarmament and development. Bulletin of the Atomic Scientists, 38(6):41–44, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thorson:1983:PDC

[Tho83] Inga Thorson. Perspective: Has disarmament a chance? Bulletin of the Atomic Scientists, 39(10):2–3, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomas:1984:CWD

[Tho84a] Urs P. Thomas. Commentary: On ‘Water dampens U.S.– Canadian relations’. Bulletin of the Atomic Scientists, 40(4):46, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Car84a].

Thornburg:1984:CEI

[Tho84b] Jonathan Thornburg. Commentary: Even if it hurts. Bulletin of the Atomic Scientists, 40(5):64, May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel84d].

Thornburg:1984:CWR

[Tho84c] Jonathan Thornburg. Commentary: Weapons from reactors. Bulletin of the Atomic Scientists, 40(7):54–55, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sub84].

Thomas:1985:DFS

[Tho85a] Claire Thomas. Defining first strike. Bulletin of the Atomic Scientists, 41(4):64, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark85h].

Thompson:1985:HD

[Tho85b] E. P. (Edward Palmer) Thompson. The Heavy Dancers. Pantheon Books, New York, NY, USA, 1985. ISBN 0-394-72895-5 (paperback), 0-394-54301-7 (jacket). xiv + 361 pp. LCCN E183.8.S65 T48 1985. US$7.95.

Thompson:1986:CES

[Tho86a] Gordon Thompson. Chernobyl: the emerging story: What happened at Reactor Four. Bulletin of the Atomic Scientists, 42(7): 26–31, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano86u].

Thorsson:1986:LSN

[Tho86b] Inga Thorsson. Letter: Swedish nuclear program. Bulletin of the Atomic Scientists, 42(7):65–66, August/September 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Joh86].

Thomas:1987:LAT

[Tho87] Claire Sherman Thomas. Letter: Amend treaty termination. Bulletin of the Atomic Scientists, 43(10):49, December 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Loe87].

Thorsson:1988:ASW

[Tho88] Inga Thorsson. After the summit: Who needs summits anyway? Bulletin of the Atomic Scientists, 44(2):36–38, March 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thorsson:1989:LB

[Tho89] Inga Thorsson. Letter: A bouquet. Bulletin of the Atomic Scientists, 45(5):46, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thompson:1990:NRT

[Tho90] Gordon Thompson. NPT review: Treaty a useful relic. Bulletin of the Atomic Scientists, 46(6):32–33, July/August 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thomas:2000:R

[Tho00] James P. Thomas. 150 requests. Bulletin of the Atomic Scientists, 56(6):40–41, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/6/40.

Thompson:2012:UTS

[Tho12] Gordon Thompson. Unmasking the truth: The science and policy of low-dose ionizing radiation. Bulletin of the Atomic Scientists, 68(3):44–50, May/June 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/68/3/44.

Thranert:1992:GBB

[Thr92] Oliver Thränert. Germans battle over blue helmets. Bulletin of the Atomic Scientists, 48(8):33–35, October 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thranert:2015:LMD

[Thr15] Oliver Thränert. Limit missile defense-or expand it? : a German response. Bulletin of the Atomic Scientists, 71(2):13–16, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Thurow:1979:CIW

[Thu79] Lester C. Thurow. Commentary: Inflation: we’re fighting yesterday’s war. Bulletin of the Atomic Scientists, 35(10):10–11, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tickner:1994:BRM

[Tic94] J. Ann Tickner. Book review: Maternal practice: Gendering War Talk, edited by Miriam Cooke and Angela Woollacott. Bulletin of the Atomic Scientists, 50(2):58–59, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tilley:1960:SP

[Til60] Paul Tilley. Scientists and politics. Bulletin of the Atomic Scientists, 16(5):187, 192, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tillet:1964:LER

[Til64] Paul Tillet. Letter to the Editor: Reflections on war, politics, and power. Bulletin of the Atomic Scientists, 20(5):28–29, May 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Timerbaev:1987:PSO

[Tim87] Roland M. Timerbaev. Perspective: A Soviet official on verification. Bulletin of the Atomic Scientists, 43(1):8–10, January/ February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [PT87].

Tinney:1967:EAN

[Tin67] E. Roy Tinney. Engineering aspects [of Nawapa]. Bulletin of the Atomic Scientists, 23(7):21–25, September 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tinsley:2023:LLB

[Tin23] Catherine H. Tinsley. Lessons learned in blood: Why we fail to use near-misses to prevent man-made disasters. Bulletin of the Atomic Scientists, 79(3):137–141, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2199565.

Tirman:1985:LMH

[Tir85] John Tirman. Letter: Militarization of high technology. Bulletin of the Atomic Scientists, 41(6):56, June/July 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sie85].

Tirman:1986:SWT

[Tir86] John Tirman. Star Wars technology threatens satellites. Bulletin of the Atomic Scientists, 42(5):28–32, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tirman:1988:BRB

[Tir88] John Tirman. Book review: Space and National Security, by Paul B. Stares. Bulletin of the Atomic Scientists, 44(1):51–52, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tiselius:1952:PSN

[Tis52] Arne Tiselius. Protest from Swedish Nobel Laureate. Bulletin of the Atomic Scientists, 8(7):256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Tismaneanu:1990:EES

[Tis90] Vladimir Tismaneanu. Eastern Europe: the story the media missed. Bulletin of the Atomic Scientists, 46(2):17–21, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tiurina:1994:BNM

[Tiu94] Tatiana Tiurina. Belarus: Neutrality, maybe. Bulletin of the Atomic Scientists, 50(1):37–40, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teeple:1966:LEP

[TK66] John B. Teeple and Arnold Kramish. Letter to the Editor: On proliferation: The danger isn’t nuclear. Bulletin of the Atomic Scientists, 22(2):32–33, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sto65b].

Tsipis:1994:LN

[TKG94] Kosta Tsipis, Michael Krepon, and Sidney N. Graybeal. Letter: It’s not necessary. Bulletin of the Atomic Scientists, 50(3):59– 60, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [GK94].

Toyoda:1979:LD

[TM79] Toshiyuki Toyoda and Alva Myrdal. Letters on disarmament. Bulletin of the Atomic Scientists, 35(3):72–80, March 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tripolskaya-Mitlyng:1994:PNN

[TM94a] Viktoria Tripolskaya-Mitlyng. Perspective: The new new Russians. Bulletin of the Atomic Scientists, 50(5):18–19, September/ October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1994:SP

[TM94b] Kosta Tsipis and Philip Morrison. An 11-step program. Bulletin of the Atomic Scientists, 50(2):41, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1994:AP

[TM94c] Kosta Tsipis and Philip Morrison. Arming for peace. Bulletin of the Atomic Scientists, 50(2):38–43, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rot94a].

Tripolskaya-Mitlyng:1996:PPT

[TM96a] Viktoria Tripolskaya-Mitlyng. Perspective: A parade of troubles. Bulletin of the Atomic Scientists, 52(6):17–19, November/ December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tripolskaya-Mitlyng:1996:PWF

[TM96b] Viktoria Tripolskaya-Mitlyng. Perspective: When a forest falls, splinters fly. Bulletin of the Atomic Scientists, 52(4):16–18, July/ August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tripolskaya-Mitlyng:1997:BEC

[TM97] Viktoria Tripolskaya-Mitlyng. Bulletin: Eastern cheers, Russian jeers, American silence. Bulletin of the Atomic Scientists, 53(1): 5–7, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1998:RAS

[TM98] Kosta Tsipis and Philip Morrison. In response: All to the sea. Bulletin of the Atomic Scientists, 54(2):48–49, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dom98].

Takagi:1990:JFP

[TN90] Junzaburo Takagi and Baku Nishio. Japan’s fake plutonium shortage. Bulletin of the Atomic Scientists, 46(8):34–38, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BP90].

Tobkin:1977:LSP

[Tob77] Donald Tobkin. Letter: The super-paranoids. Bulletin of the Atomic Scientists, 33(7):2, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav77a].

Tobey:2012:NSA

[Tob12] William Tobey. Nuclear scientists as assassination targets. Bulletin of the Atomic Scientists, 68(1):61–69, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/68/1/61.

Todd:1955:LEM

[Tod55] David Todd. Letter to the Editor: [Madame Joliot and the American Chemical Society]. Bulletin of the Atomic Scientists, 11(10): 380, December 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hai54, Eme54].

Todd:1964:ICO

[Tod64a] Murray Todd. International cooperation: “Our common enterprise”: A non-apocalyptic view. Bulletin of the Atomic Scientists, 20(2):27–29, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1964:RSS

[Tod64b] W. Murray Todd. Reports: Science in the State Department. Bulletin of the Atomic Scientists, 20(10):27–29, December 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1978:BRW

[Tod78a] W. Murray Todd. Book reviews: The Worldwatch Institute and its books, papers and themes. Bulletin of the Atomic Scientists, 34(9):45–48, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1978:BBT

[Tod78b] W. Murray Todd. Books: Science and Technology and American Diplomacy: An Extended Study of the Interactions of Science and Technology with U.S. Foreign Policy. Bulletin of the Atomic Scientists, 34(6):53–55, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1979:BRB

[Tod79] W. Murray Todd. Book reviews: U.S. Food: Making the Most of a Global Resource, by David W. McClintock; Worldwatch Paper 29: Resource Trends and Population Policy: A Time for Reassessment, by Lester R. Brown; U.S. International Population Policy: Third Annual Report of the NSC Ad Hoc Group on Population Policy. Bulletin of the Atomic Scientists, 35(10):38–41, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1980:BRB

[Tod80] W. Murray Todd. Book review: Faith, Science and the Future, edited by Paul Abrecht. Bulletin of the Atomic Scientists, 36(2): 48, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1981:BRB

[Tod81] W. Murray Todd. Book reviews: The Wooing of Earth: New Perspectives on Man’s Use of Nature, by René Dubos; Quest — Reflections on Medicine, Science, and Humanity, by René Dubos and Jean-Paul Escande, translated by Patricia Ranum. Bulletin of the Atomic Scientists, 37(5):44–45, May 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Toll:1958:SUL

[Tol58] John S. Toll. Scientists urge lifting travel restrictions. Bulletin of the Atomic Scientists, 14(8):326–328, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tomori:2014:WBA

[Tom14] Oyewale Tomori. Winning the battle against emerging pathogens: a Nigerian response. Bulletin of the Atomic Scientists, 70(4): 14–17, July 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Topchiev:1956:AMC

[Top56] A. V. Topchiev. [Address to Milan conference, August 1955]. Bulletin of the Atomic Scientists, 12(2):44–45, February 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Topchev:1958:CPE

[Top58a] A. V. Topchev. Comments on Pugwash: From the East: To prevent atomic war. Bulletin of the Atomic Scientists, 14(3):118– 120, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Topchiev:1958:TPC

[Top58b] A. V. Topchiev. Third Pugwash Conference: Disarmament and international tension. Bulletin of the Atomic Scientists, 14(10): 405–408, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Topchiev:1962:PSC

[Top62] A. V. Topchiev. Prospects for scientific cooperation. Bulletin of the Atomic Scientists, 18(1):16–20, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Topchiev:1963:ISS

[Top63] A. V. Topchiev. Interdependence of science and society. Bulletin of the Atomic Scientists, 19(3):7, 9–11, March 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Torstenson:1961:BBF

[Tor61] Joel S. Torstenson. Books: The Future of Mankind, by Karl Jaspers. Bulletin of the Atomic Scientists, 17(5–6):245, May/ June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Torrie:1977:LJG

[Tor77a] Ralph Torrie. Letter: Jumping the gun. Bulletin of the Atomic Scientists, 33(5):7, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Maw77].

Torrie:1977:LWR

[Tor77b] Ralph D. Torrie. Letter: From wheels to reactors. Bulletin of the Atomic Scientists, 33(2):2, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Flo76b].

Torry:1977:LBW

[Tor77c] Volta Torry. Letter: Beautiful windpower. Bulletin of the Atomic Scientists, 33(1):6, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sør76].

Torres:2019:PRA

[Tor19] Phil Torres. The possibility and risks of artificial general intelligence. Bulletin of the Atomic Scientists, 75(3):105–108, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1604873.

Toschi:1982:CDP

[Tos82] James P. Toschi. Commentary: Defense policies and authoritarianism. Bulletin of the Atomic Scientists, 38(6):70–72, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Totter:1970:LT

[Tot70] John R. Totter. Letter to Tamplin. Bulletin of the Atomic Scientists, 26(7):48, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [TG70].

Tovish:1989:LDK

[Tov89] Aaron Tovish. Letter: Don’t knock test ban conference. Bulletin of the Atomic Scientists, 45(8):44–45, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cle89].

Townsend:1966:IPS

[Tow66] James R. Townsend. Internal politics since 1956. Bulletin of the Atomic Scientists, 22(6):58–65, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Toyoda:1982:HY

[Toy82] Toshiyuki Toyoda. Hideki Yukawa 1907–1981. Bulletin of the Atomic Scientists, 38(8):65, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Toyoda:1984:PSJ

[Toy84a] Toshiyuki Toyoda. Peace and security in Japan. Bulletin of the Atomic Scientists, 40(2):15, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Toyoda:1984:SLP

[Toy84b] Toshiyuki Toyoda. Scientists look at peace and security. Bulletin of the Atomic Scientists, 40(2):16–19, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Toyoda:1985:JPS

[Toy85] Toshiyuki Toyoda. Japan’s policies since 1945. Bulletin of the Atomic Scientists, 41(7):57–62, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Toyoda:1995:LIE

[Toy95] Toshiyuki Toyoda. Letter: Indefinite extension, no. Bulletin of the Atomic Scientists, 51(3):68, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See opposing view [BT95].

Toyoda:1996:LXF

[Toy96] Toshiyuki Toyoda. Letter: The “X” factor. Bulletin of the Atomic Scientists, 52(1):61, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo95o].

Tracy:2021:OUR

[TPPB21] Cameron L. Tracy, Sulgiye Park, Mariia Plevaka, and Ekaterina Bogdanova. Opportunities for US–Russian collaboration on the safe disposal of nuclear waste. Bulletin of the Atomic Scientists, 77(3):146–152, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1912276.

Tribe:1980:SRB

[TR80] Laurence H. Tribe and David H. Remes. Some reflections on The Progressive Case: publish and perish? Bulletin of the Atomic Scientists, 36(3):20–24, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [STRS80, O’B80].

Trapp:1951:E

[Tra51] Joseph Trapp. Epigram. Bulletin of the Atomic Scientists, 7(11): 324, November 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Trager:1971:BBR

[Tra71] Frank N. Trager. Books: China: The Revolution Continued, by Jan Myrdal and Gun Kessle; translated from the revised Swedish edition by Paul Britten Austin. Bulletin of the Atomic Scientists, 27(7):54, September 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Train:1972:ROP

[Tra72a] Russell E. Train. Response [to Olof Palme]. Bulletin of the Atomic Scientists, 28(7):45, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pal72].

Train:1972:RT

[Tra72b] Russell E. Train. Response [to Tang Ke]. Bulletin of the Atomic Scientists, 28(7):55, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ke72].

Train:1972:USP

[Tra72c] Russell E. Train. The United States position. Bulletin of the Atomic Scientists, 28(7):43–44, September 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Train:1973:EPE

[Tra73] Russell E. Train. Energy problems and environmental concern. Bulletin of the Atomic Scientists, 29(9):43–47, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Por74].

Trachtenberg:1990:BRW

[Tra90] Marc Trachtenberg. Book review: Who cares about flexible response? Guardians of the Arsenal: The Politics of Nuclear Strategy, by Janne Nolan. Bulletin of the Atomic Scientists, 46(4): 48–49, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tracy:2024:SRP

[Tra24] Cameron L. Tracy. Sociotechnical risks posed by the geologic disposal of weapons plutonium. Bulletin of the Atomic Scientists, 81(1):36–42, 2024. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2024.2441045.

Treder:2007:NW

[Tre07] Mike Treder. Nanotechnology at war. Bulletin of the Atomic Scientists, 63(6):57–58, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/63/6/57.

Treuer:2018:PMS

[Tre18] Galen A. Treuer. The psychology of Miami’s struggle to adapt to sea-level rise. Bulletin of the Atomic Scientists, 74(3):155– 159, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1461970.

Trenin:2019:RVU

[Tre19] Dmitri Trenin. Russian views of US nuclear modernization. Bulletin of the Atomic Scientists, 75(1):14–18, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2019.1555991.

Trevan:2021:PUC

[Tre21] Tim Trevan. A perspective on UNSCOM culture. Bulletin of the Atomic Scientists, 77(4):195–198, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1943873.

Tribus:1971:ITS

[Tri71] Myron Tribus. II. Technology and society — the real issues. Bulletin of the Atomic Scientists, 27(10):27–30, December 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Shu72b] and reply [Tri72].

Tribus:1972:CR

[Tri72] Myron Tribus. Communications: a reply. Bulletin of the Atomic Scientists, 28(5):3–4, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tri71, Shu72b, Shu72a].

Trimble:1977:LPD

[Tri77] M. H. Trimble. Letter: Photovoltaic devices. Bulletin of the Atomic Scientists, 33(1):4, 6, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wol76].

Troitskaya:1958:IGY

[Tro58] V. A. Troitskaya. International Geophysical Year activities in the Soviet Union. Bulletin of the Atomic Scientists, 14(5):173–176, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Trofimenko:1979:MSI

[Tro79] Henry Trofimenko. From Moscow: SALT II: a fair bargain. Bulletin of the Atomic Scientists, 35(6):30–34, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Troitskiy:2020:WUR

[Tro20] Mikhail Troitskiy. Why US–Russian arms control can succeed even in a climate of confrontation. Bulletin of the Atomic Scientists, 76(2):91–96, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1728978.

Truman:1946:CNP

[Tru46] Harry S. Truman. Congressional news: President Truman endorses the McMahon Bill. Bulletin of the Atomic Scientists of Chicago, 1(5):5, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Truman:1947:TAR

[Tru47] Harry S. Truman. Truman accepts responsibility for bomb. Bulletin of the Atomic Scientists, 3(2):40, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Truman:1948:PTA

[Tru48a] Harry S. Truman. President Truman answers Dewey. Bulletin of the Atomic Scientists, 4(11):336, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dew48].

Truman:1948:PTS

[Tru48b] Harry S. Truman. President Truman speaks to the scientists. Bulletin of the Atomic Scientists, 4(10):291–293, October 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Based on address of 13 September 1948 to the American Association for the Advancement of Science in Washington, DC.

Truman:1949:PTB

[Tru49] Harry S. Truman. President Truman backs AEC. Bulletin of the Atomic Scientists, 5(6–7):185, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Truman:1950:TSN

[Tru50] Harry S. Truman. Truman’s statement on National Science Foundation Bill. Bulletin of the Atomic Scientists, 6(6):165, June 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Truman:1951:NDP

[Tru51] Harry S. Truman. National dispersal program forecast by presidential statement. Bulletin of the Atomic Scientists, 7(9):263, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Truman:1953:PTF

[Tru53] Harry S. Truman. President Truman’s farewell address. Bulletin of the Atomic Scientists, 9(1):4–6, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Trudeau:1983:PCP

[Tru83] Pierre Trudeau. Perspective: The Canadian Prime Minister speaks. Bulletin of the Atomic Scientists, 39(8):2–3, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Tsi83f, RI83].

Trudeau:1984:API

[Tru84] Pierre Elliott Trudeau. Agenda 1984: A peace initiative from Canada. Bulletin of the Atomic Scientists, 40(1):15–16, 18–19, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Trudeau:1985:WLM

[Tru85] Pierre Elliott Trudeau. World leaders must reassert primacy. Bulletin of the Atomic Scientists, 41(2):10–13, February 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Smi85b].

Thompson:1992:PLX

[TS92] Gordon Thompson and Steven C. Sholly. Perspective: Let’s Xout the K. Bulletin of the Atomic Scientists, 48(2):14–15, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tucker:1999:UT

[TS99] Jonathan B. Tucker and Amy Sands. An unlikely threat. Bulletin of the Atomic Scientists, 55(4):46–52, July/August 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/4/46.

Teller:2001:MTC

[TS01] Edward Teller and Judith L. Shoolery. Memoirs: a Twentieth-Century Journey in Science and Politics. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0532-X, 1-903985-12-9. xii + 628 pp. LCCN QC16.T37 M55 2001. URL http:/

/alsos.wlu.edu/information.aspx?id=784;   http://www.
loc.gov/catdir/enhancements/fy0833/2001097880-d.html;
 http://www.loc.gov/catdir/toc/fy02/2001097880.html.

Tavares:2022:YGC

[TS22] Joseph Tavares and Kori Schake. Not your grandparents’ Cold War: Why America should emphasize economic rather than military strategies in its rivalry with China. Bulletin of the Atomic Scientists, 78(6):324–328, 2022. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2022.2132739.

Tschudin:1987:LSM

[Tsc87] Patrik Tschudin. Letter: Swiss militarism. Bulletin of the Atomic Scientists, 43(5):53–54, June 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tay87].

Tsipis:1972:BBN

[Tsi72a] Kosta Tsipis. Books: Impact of New Technologies on the Arms Race, edited by B. T. Feld, T. Greenwood, G. W. Rathjens, and S. Weinberg. Bulletin of the Atomic Scientists, 28(1):47–48, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1972:HBM

[Tsi72b] Kosta Tsipis. Hiding behind the military–industrial complex. Bulletin of the Atomic Scientists, 28(6):20–23, June 1972. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cae73].

Tsipis:1975:DMM

[Tsi75a] Kosta Tsipis. Digital map making: How an intercontinental shot can hit within 10 meters. Bulletin of the Atomic Scientists, 31(4): 22–23, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1975:LRC

[Tsi75b] Kosta Tsipis. The long-range cruise missile. Bulletin of the Atomic Scientists, 31(4):14–26, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1976:SBS

[Tsi76] Kosta Tsipis. Security blankets of the superpowers. Bulletin of the Atomic Scientists, 32(9):19–20, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1977:ASR

[Tsi77a] Kosta Tsipis. The armed services’ role. Bulletin of the Atomic Scientists, 33(6):30–31, June 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1977:BBW

[Tsi77b] Kosta Tsipis. The building blocks of weapons development. Bulletin of the Atomic Scientists, 33(4):41–42, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1977:IMR

[Tsi77c] Kosta Tsipis. [Introduction: military R&D]. Bulletin of the Atomic Scientists, 33(2):32, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1977:SM

[Tsi77d] Kosta Tsipis. Science and the military. Bulletin of the Atomic Scientists, 33(1):10–11, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1978:EFT

[Tsi78a] Kosta Tsipis. Editor’s foreword: Time for action. Bulletin of the Atomic Scientists, 34(2):33, February 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1978:NLI

[Tsi78b] Kosta Tsipis. A neo-Luddite’s interpretation. Bulletin of the Atomic Scientists, 34(5):42, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1978:PI

[Tsi78c] Kosta Tsipis. Pragmatic insights. Bulletin of the Atomic Scientists, 34(10):40, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1980:SWP

[Tsi80] Kosta Tsipis. Scientists and weapons procurement. Bulletin of the Atomic Scientists, 36(6):41–43, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1982:BP

[Tsi82a] Kosta Tsipis. Breaking with the past. Bulletin of the Atomic Scientists, 38(5):4–5, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1982:EWE

[Tsi82b] Kosta Tsipis. Extreme wrong on the extreme right. Bulletin of the Atomic Scientists, 38(4):3–5, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BT82].

Tsipis:1982:PCC

[Tsi82c] Kosta Tsipis. Perspective: Countering counterforce. Bulletin of the Atomic Scientists, 38(8):3–4, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1982:TSY

[Tsi82d] Kosta Tsipis. Thirty-seven years of nuclear weapons: An introduction. Bulletin of the Atomic Scientists, 38(10):46, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1983:CDF

[Tsi83a] Kosta Tsipis. Commentary: Downing of flight 007. Bulletin of the Atomic Scientists, 39(9):63–64, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dre84b].

Tsipis:1983:CSB

[Tsi83b] Kosta Tsipis. Commentary: Not such a bargain after all. Bulletin of the Atomic Scientists, 39(9):54–55, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Asp83] and comments [Sch84a, You84a].

Tsipis:1983:EIB

[Tsi83c] Kosta Tsipis. [Editor’s introduction to The spiral of peril: a narrative of the nuclear arms race]. Bulletin of the Atomic Scientists, 39(1):10–11, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1983:NWTa

[Tsi83d] Kosta Tsipis. [Note on The weapons tutorial — Part 3: Research, development and production]. Bulletin of the Atomic Scientists, 39(3):45, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1983:NWTb

[Tsi83e] Kosta Tsipis. [Note on The weapons tutorial — Part 4: Nuclear weapons in Europe]. Bulletin of the Atomic Scientists, 39(4):17, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1983:PAR

[Tsi83f] Kosta Tsipis. Perspective: An American responds. Bulletin of the Atomic Scientists, 39(8):3–5, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tru83].

Tsipis:1983:SENb

[Tsi83g] Kosta Tsipis. [Special Editor’s note on “The gaps of wrath”]. Bulletin of the Atomic Scientists, 39(8):28, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1983:SENa

[Tsi83h] Kosta Tsipis. [Special Editor’s note on “When the bomb falls”]. Bulletin of the Atomic Scientists, 39(6):40–47, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1983:WTP

[Tsi83i] Kosta Tsipis. The weapons tutorial — Part 2: Inside the mushroom cloud. Bulletin of the Atomic Scientists, 39(2):23–27, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Parts 1 [Mor83c], 3 [Mar83b], 4 [Hol83c], 5 [Sar83], and 6 [Rom83], and erratum [Ano83m].

Tsipis:1984:CMS

[Tsi84a] Kosta Tsipis. Commentary: Max Stanley. Bulletin of the Atomic Scientists, 40(10):32, December 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1984:SEN

[Tsi84b] Kosta Tsipis. [Special editor’s note on “A history of strategic arms limitations”]. Bulletin of the Atomic Scientists, 40(1):30, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1987:PUG

[Tsi87] Kosta Tsipis. Perspective: The uncommon ground of security debate. Bulletin of the Atomic Scientists, 43(6):10–11, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cat87].

Tsipis:1988:CMS

[Tsi88] Kosta Tsipis. Cruise missiles should not stop START. Bulletin of the Atomic Scientists, 44(9):36–40, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1989:BRB

[Tsi89a] Kosta Tsipis. Book review: Flanagan’s Version, by Dennis Flanagan. Bulletin of the Atomic Scientists, 45(2):44–45, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1989:LSM

[Tsi89b] Kosta Tsipis. Letter: Subs, not missiles. Bulletin of the Atomic Scientists, 45(1):66, January/February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kre88f].

Tsipis:1989:PAC

[Tsi89c] Kosta Tsipis. Perspective: After the Cold War: New tasks for arms controllers. Bulletin of the Atomic Scientists, 45(6):7–8, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsipis:1990:PTR

[Tsi90] Kosta Tsipis. Perspective: Time for rebirth of civilian R&D. Bulletin of the Atomic Scientists, 46(9):11–12, November 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tsypkin:1983:SUC

[Tsy83] Mikhail Tsypkin. The Soviet Union: The conscripts. Bulletin of the Atomic Scientists, 39(5):28–32, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Taubenfeld:1969:RII

[TT69] Rita F. Taubenfeld and Howard J. Taubenfeld. Reports: The international implications of weather modification. Bulletin of the Atomic Scientists, 25(1):43–45, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tovish:1983:BND

[TT83] Aaron Tovish and Nina Tovish. Breaking the nuclear deadlock: Neutral nations peace initiative. Bulletin of the Atomic Scientists, 39(8):58–59, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tuck:1962:LEK

[Tuc62] James L. Tuck. Letter to the Editor: Kapitza. Bulletin of the Atomic Scientists, 18(7):37, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap62].

Tucker:1964:LEB

[Tuc64] Frank H. Tucker. Letter to the Editor: Berlin. Bulletin of the Atomic Scientists, 20(10):22, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul64b].

Tucker:1966:PSA

[Tuc66] Robert C. Tucker. Proliferation and Soviet–American relations. Bulletin of the Atomic Scientists, 22(8):14–18, October 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tuck:1972:BBI

[Tuc72] James L. Tuck. Books: Beyond the Ivory Tower: The Frontiers of Public and Private Science, by Sir Solly Zuckerman. Bulletin of the Atomic Scientists, 28(1):46, January 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tucker:1977:BWB

[Tuc77] Anthony Tucker. Books: ‘whistle-blowing’ in Britain: Superstar technologies. Bulletin of the Atomic Scientists, 33(5):61–62, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tucker:1997:BRC

[Tuc97] Jonathan B. Tucker. Book review: Chemistry lesson: The Eleventh Plague: The Politics of Biological and Chemical Warfare, by Leonard A. Cole, W. H. Freeman and Co., 1996, 280 pages, $22.95. Bulletin of the Atomic Scientists, 53(2):57–58, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tucker:2001:CWB

[Tuc01] Jonathan B. Tucker. Chemical weapons: Buried in the backyard. Bulletin of the Atomic Scientists, 57(5):51–56, September/ October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/57/
5/51.

Tucker:2008:BOB

[Tuc08] Jonathan B. Tucker. The body’s own bioweapons. Bulletin of the Atomic Scientists, 64(1):16–22, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/1/16.

Tucker:2010:BRL

[Tuc10a] Jonathan B. Tucker. A biosecurity reading list. Bulletin of the Atomic Scientists, 66(1):63–73, January 2010. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/66/1/63.

Tucker:2010:CBC

[Tuc10b] Jonathan B. Tucker. The convergence of biology and chemistry: Implications for arms control verification. Bulletin of the Atomic Scientists, 66(6):56–66, November/December 2010. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/66/6/56.

Tudor:2021:HCC

[Tud21] Ioan Tudor. How countries can build on UNSCOM’s legacy to solve today’s problems. Bulletin of the Atomic Scientists, 77 (4):192–194, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1941590.

Tufts:1954:LEM

[Tuf54] Merrill Tufts. Letter to the Editor: March 1954 editorial in the Bulletin. Bulletin of the Atomic Scientists, 10(6):212, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tugwell:1951:CK

[Tug51] Rexford G. Tugwell. The consequences of Korea. Bulletin of the Atomic Scientists, 7(5):133–138, 148, May 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tumarkin:1987:BRB

[Tum87] Nina Tumarkin. Book review: Shadows and Whispers: Power Politics inside the Kremlin from Brezhnev to Gorbachev, by Dusko Doder. Bulletin of the Atomic Scientists, 43(6):57–58, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turkevich:1950:ASS

[Tur50] Anthony Turkevich. AEC seventh semiannual report. Bulletin of the Atomic Scientists, 6(5):153–160, May 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turkevich:1951:PAM

[Tur51] Anthony Turkevich. Progress and activities in major atomic energy programs in 1950: Ninth semiannual report of the USAEC: Part one. Bulletin of the Atomic Scientists, 7(6):177–182, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See also Part II [Bos51].

Turkevich:1952:MAU

[Tur52] Anthony Turkevich. Major activities in U.S. atomic energy programs, January–June, 1951: Summary of the tenth AEC semiannual report. Bulletin of the Atomic Scientists, 8(1):23–26, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turkevich:1953:APS

[Tur53] Anthony Turkevich. Assuring public safety in continental weapons tests: AEC thirteenth semiannual report. Bulletin of the Atomic Scientists, 9(3):85–89, April 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turner:1954:KTC

[Tur54] Louis Turner. Karl Taylor Compton: an appreciation. Bulletin of the Atomic Scientists, 10(7):296, 304, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turkevich:1960:GTB

[Tur60] Anthony Turkevich. Geneva test ban negotiations: USSR, U.K., and U.S. reports. Bulletin of the Atomic Scientists, 16(2):38, February 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turkevich:1973:MEP

[Tur73] Anthony L. Turkevich. A method of exploring planets: The chemical analysis of the lunar surface on Surveyor. Bulletin of the Atomic Scientists, 29(10):27–34, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turk:1976:LAI

[Tur76] Amos Turk. Letter: Anti-Israel. Bulletin of the Atomic Scientists, 32(9):7, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Turkisher:1977:MUT

[Tur77] K. Ch. Turkisher. Monologue on an unpopular theme: some reflections on nuclear vengeance. Bulletin of the Atomic Scientists, 33(2):48–51, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turchin:1978:BSU

[Tur78] Valentin Turchin. Boycotting the Soviet Union. Bulletin of the Atomic Scientists, 34(7):7–11, September 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turner:1987:LCC

[Tur87] Paul Turner. Letter: Chernobyl containment. Bulletin of the Atomic Scientists, 43(2):64, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turner:1998:PCS

[Tur98] Stansfield Turner. Perspective: The case for strategic escrow. Bulletin of the Atomic Scientists, 54(2):16–17, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turpen:2013:RNS

[Tur13] Libby Turpen. Revitalizing the national security labs: Beyond the nuclear deterrent. Bulletin of the Atomic Scientists, 69(1): 53–61, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/69/1/53.

Turrentine:2014:CBC

[Tur14] Tom Turrentine. California: Beyond cars? Bulletin of the Atomic Scientists, 70(5):54–61, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Turco:2017:NFS

[Tur17] Richard Turco. Nuclear foreboding: Shadows cast by nuclear winter. Bulletin of the Atomic Scientists, 73(4):240–243, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1338017.

Turner:2018:NWD

[Tur18] Katlyn Turner. New ways to detect nuclear misbehavior. Bulletin of the Atomic Scientists, 74(1):2–6, 2018. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1413054.

Tuve:1953:TNR

[Tuv53] Merle A. Tuve. Technology and national research policy. Bulletin of the Atomic Scientists, 9(8):290–293, October 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Todd:1971:CAN

[TV71] W. M. Todd and John Voss. The consortium of academies: A new way to found international scholarly institutions. Bulletin of the Atomic Scientists, 27(2):29–32, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1983:FH

[TvH83] Edward Teller and Frank von Hippel. On facts and hopes. Bulletin of the Atomic Scientists, 39(4):42–44, April 1983. CODEN BASIAP. ISSN 0096-3402. Response to [vH83a].

Takubo:2017:FYI

[TvH17] Masafumi Takubo and Frank von Hippel. Forty years of impasse: The United States, Japan, and the plutonium problem. Bulletin of the Atomic Scientists, 73(5):337–343, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1364007.

Thornton:1966:LEM

[TW66] Richard C. Thornton and James West. Letter to the Editor: The map of China. Bulletin of the Atomic Scientists, 22(9):24– 25, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Twitchell:2023:LGL

[Twi23] Jeremy Twitchell. Laying the groundwork for long-duration energy storage. Bulletin of the Atomic Scientists, 79(6):372–376, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2266939.

Taylor:1961:MRC

[TWSM61] Wm. Palmer Taylor, A. H. Whitelaw, J. David Singer, and Arthur W. Munk. More replies on China. Bulletin of the Atomic Scientists, 17(10):430, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Tomonaga:1975:BND

[TY75] Sin-Itiro Tomonaga and Hideki Yikawa. Beyond nuclear deterrence. Bulletin of the Atomic Scientists, 31(10):9, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Presented at a symposium on “A New Design toward Complete Nuclear Disarmament” held in Kyoto, Japan, 28 August–1 September 1975.

Tyler:1989:LUI

[Tyl89] John Tyler. Letter: Uranium imports: so what? Bulletin of the Atomic Scientists, 45(10):50, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lan89b].

Tyner:1975:LRS

[Tyn75] Mrs. Mack Tyner, Jr. Letters: Runaway stork. Bulletin of the Atomic Scientists, 31(2):3, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mul74b].

Uppman:1962:LEH

[UC62] Ingerard Uppman and Alice Cooper. Letter to the Editor: Hiroshima. Bulletin of the Atomic Scientists, 18(5):41, May 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Usborne:1956:WPW

[UH56] Henry Usborne and P. E. Hodgson. World politics and world government. Bulletin of the Atomic Scientists, 12(4):125–131, April 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1995:BRN

[Uhl95a] Walter C. Uhler. Book review: No thanks to Reagan: The Great Transition: American–Soviet Relations and the End of the Cold War, by Raymond L. Garthoff. Bulletin of the Atomic Scientists, 51(3):65–67, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1995:BRU

[Uhl95b] Walter C. Uhler. Book review: Under the gun: In the Shadow of War: The United States Since the 1930s, by Michael S. Sherry. Bulletin of the Atomic Scientists, 51(6):68–71, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1996:BRC

[Uhl96] Walter C. Uhler. Book review: Combating conversion: A Farewell to Arms? Russia’s Struggles with Defense Conversion, by Kevin P. O’Prey. Bulletin of the Atomic Scientists, 52(4):57–59, July/ August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1997:BRB

[Uhl97a] Walter C. Uhler. Book review: Defense Addiction: Can America Kick the Habit?, by Sanford Gottlieb. Bulletin of the Atomic Scientists, 53(6):60–61, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1997:BRE

[Uhl97b] Walter C. Uhler. Book review: Ending the Cold War: The Gorbachev, by Archie Brown. Bulletin of the Atomic Scientists, 53(1): 65–66, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1998:BRB

[Uhl98a] Walter C. Uhler. Book review: Spoils of War: The Human Cost of America’s Arms Trade, by John Tirman. Bulletin of the Atomic Scientists, 54(2):64–65, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:1998:BRD

[Uhl98b] Walter C. Uhler. Book review: Debunking the “red scare”: China’s Security: The New Roles of the Military, by Mel Gurtov and Byong-Moo Hwang. Bulletin of the Atomic Scientists, 54 (6):58–59, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Uhler:2000:MS

[Uhl00] Walter C. Uhler. Misreading the Soviets. Bulletin of the Atomic Scientists, 56(4):68–69, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/4/68.

Uhler:2001:UNM

[Uhl01a] Walter C. Uhler. Unilateral no more? Bulletin of the Atomic Scientists, 57(6):70–71, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/57/6/70.

Uhler:2001:VT

[Uhl01b] Walter C. Uhler. The vision thing. Bulletin of the Atomic Scientists, 57(3):70–72, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/3/70.

Uhler:2003:TWI

[Uhl03a] Walter C. Uhler. Triumphalism, or the “Wolfowitz Indiscretion”. Bulletin of the Atomic Scientists, 59(2):72–73, March 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/2/72.

Uhler:2003:UCE

[Uhl03b] Walter C. Uhler. Undone by current events. Bulletin of the Atomic Scientists, 59(4):66–67, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/4/66.

Uhler:2004:DD

[Uhl04a] Walter C. Uhler. Democracy or dominion? Bulletin of the Atomic Scientists, 60(1):52–58, January 2004. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/1/52.

Uhler:2004:PT

[Uhl04b] Walter C. Uhler. Preempting the truth. Bulletin of the Atomic Scientists, 60(5):63–66, September/October 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/5/63.

Uhler:2005:EE

[Uhl05] Walter C. Uhler. Engage or enrage. Bulletin of the Atomic Scientists, 61(5):65–67, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/61/5/65.

Ulgen:2015:ZRT

[Ulg15] Sinan Ulgen. Is “zero” the right target for disarmament?: a Turkish response. Bulletin of the Atomic Scientists, 71(1):95–97, January 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ullmann:1963:BBW

[Ull63] John E. Ullmann. Books: The Weapons Acquisition Process, by Merton J. Peck and Frederic M. Scherer. Bulletin of the Atomic Scientists, 19(6):27–28, June 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1967:CLS

[UM67] Harold C. Urey and Sandra B. McPherson. Comments: Lorenz and the social responsibility of the scientific community. Bulletin of the Atomic Scientists, 23(4):35–36, April 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab66b].

Umpleby:1975:LWP

[Ump75] Stuart A. Umpleby. Letter: The wrong point. Bulletin of the Atomic Scientists, 31(9):3, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ungerer:1978:EI

[Ung78a] Werner Ungerer. Energy islands. Bulletin of the Atomic Scientists, 34(8):52, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ungerer:1978:LDN

[Ung78b] Werner Ungerer. Letter: Denial is no solution. Bulletin of the Atomic Scientists, 34(4):2, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Flo78].

Unger:1983:YCI

[Ung83] Stephen H. Unger. You can’t have it both ways. Bulletin of the Atomic Scientists, 39(7):17–18, August/September 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Originally presented at the annual meeting of the American Association for the Advancement of Science in Denver in May 1983: see addendum [Ano83b].

Ungar:1985:AMM

[Ung85] Sanford J. Ungar. Africa: The military money drain. Bulletin of the Atomic Scientists, 41(8):31–34, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

USAEC:1947:RDD

[Uni47] United States Atomic Energy Commission. Report on domestic distribution of radioisotopes. Bulletin of the Atomic Scientists, 3 (10):305–308, October 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

UIP:1954:LSJ

[Uni54] University of Illinois Physicists. [Letter of support for J. Robert Oppenheimer]. Bulletin of the Atomic Scientists, 10(5):191, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Letter signed by 27 individuals.

UNSCEAR:1957:RMP

[Uni57] United Nations Scientific Committee on the Effects of Atomic Radiation. The responsibilities of the medical profession in the use of X-rays and other ionizing radiation. Bulletin of the Atomic Scientists, 13(4):137–138, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

UCS:1978:DNA

[Uni78] Union of Concerned Scientists. Declaration on the nuclear arms race: an appeal to the President and Congress to take the initiative in controlling strategic weapons technology. Bulletin of the Atomic Scientists, 34(3):8–10, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Upton:1963:BBU

[Upt63] Arthur C. Upton. Books: Report of the United Nations Scientific Committee on the Effects of Atomic Radiation. Bulletin of the Atomic Scientists, 19(4):32–33, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urayama:2001:MDJ

[Ura01] Kori Urayama. Missile defense: Japan’s wait-and-see approach. Bulletin of the Atomic Scientists, 57(6):33–35, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/6/33.

Urey:1945:PHA

[Ure45] H. C. Urey. Pearl Harbor anniversary and the Moscow conference. Bulletin of the Atomic Scientists of Chicago, 1(1):1, December 10, 1945. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Statement released on 7 December 1945, the fourth anniversary of the Japanese military’s aeriel attack on Hawaii.

Urey:1946:SVW

[Ure46a] H. C. Urey. A scientist views the world situation. Bulletin of the Atomic Scientists of Chicago, 1(5):4, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1946:AEW

[Ure46b] Harold C. Urey. Atomic energy and world peace. Bulletin of the Atomic Scientists, 2(9–10):2–4, November 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1946:HCU

[Ure46c] Harold C. Urey. H. C. Urey on State Dept. report. Bulletin of the Atomic Scientists, 1(8):13, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1947:ACC

[Ure47] Harold C. Urey. An alternative course for the control of atomic energy. Bulletin of the Atomic Scientists, 3(6):139–142, 168, June 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1948:AEC

[Ure48a] Harold C. Urey. Atomic energy control is impossible without world government. Bulletin of the Atomic Scientists, 4(12):365– 366, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fri48a].

Urey:1948:CDB

[Ure48b] Harold C. Urey. Comments on Dr. Bridgman’s article: Scientists should assume responsibility. Bulletin of the Atomic Scientists, 4(3):72–73, March 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bri48a].

Urey:1948:FDR

[Ure48c] Harold C. Urey. Freedom demands responsibility — the scientist’s role in society: A scientists’ strike will not insure peace. Bulletin of the Atomic Scientists, 4(11):337, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a].

Urey:1948:WMP

[Ure48d] Harold C. Urey. We must protect the last increment of ability. Bulletin of the Atomic Scientists, 4(6):175, June 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1949:AER

[Ure49a] Harold C. Urey. The atomic explosion in Russia: Needed: Less witch hunting and more work. Bulletin of the Atomic Scientists, 5(10):265, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1949:PP

[Ure49b] Harold C. Urey. The paramount problem of 1949. Bulletin of the Atomic Scientists, 5(10):283–288, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1950:SAB

[Ure50] Harold C. Urey. Should America build the H-bomb? Bulletin of the Atomic Scientists, 6(3):72–73, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1954:I

[Ure54] Harold C. Urey. Issues of 1954. Bulletin of the Atomic Scientists, 10(1):12, 32, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1955:CP

[Ure55] Harold C. Urey. “Chemically pure”. Bulletin of the Atomic Scientists, 11(2):65, 72, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1956:LEH

[Ure56] Harold C. Urey. Letter to the Editor: On Harold Green’s review of William A. Reuben’s book The Atom Spy Hoax. Bulletin of the Atomic Scientists, 12(1):31, January 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1961:LER

[Ure61] Harold Urey. Letter to the Editor: The real problem in policing a test ban. Bulletin of the Atomic Scientists, 17(2):40, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1967:CAS

[Ure67] Harold C. Urey. Comments: Affording the space program. Bulletin of the Atomic Scientists, 23(2):24–25, February 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ano67e, Bor68].

Urey:1969:OHM

[Ure69a] Harold C. Urey. Origin and history of the moon. Bulletin of the Atomic Scientists, 25(7):46–51, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1969:SPP

[Ure69b] Harold C. Urey. The space program and problems of the origin of the Moon. Bulletin of the Atomic Scientists, 25(4):24–26, 30, April 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urey:1973:MO

[Ure73] Harold C. Urey. The Moon and its origin. Bulletin of the Atomic Scientists, 29(9):5–10, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urigashvili:1994:GDW

[Uri94a] Besik Urigashvili. Georgia: Damn this war …. Bulletin of the Atomic Scientists, 50(1):20–24, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urigashvili:1994:SCF

[Uri94b] Besik Urigashvili. A sudden change of fortune. Bulletin of the Atomic Scientists, 50(1):22, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Urigashvili:1994:TBT

[Uri94c] Besik Urigashvili. The Transcaucasus: Blood ties. Bulletin of the Atomic Scientists, 50(1):18–19, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Usborne:1947:CWG

[Usb47] Henry C. Usborne. The crusade for world government. Bulletin of the Atomic Scientists, 3(12):359–360, December 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Usborne:1952:FE

[Usb52] Henry Usborne. The federalists in Europe. Bulletin of the Atomic Scientists, 8(5):157–158, 165, June 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Usmani:1979:EBS

[Usm79] I. H. Usmani. Energy banks for small villages. Bulletin of the Atomic Scientists, 35(6):40–44, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ustiugov:1993:BOM

[Ust93a] Mikhail Ustiugov. Big Oil moves in. Bulletin of the Atomic Scientists, 49(8):44–47, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ustiugov:1993:EW

[Ust93b] Mikhail Ustiugov. An embarrassment of weapons. Bulletin of the Atomic Scientists, 49(8):48–50, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ustiugov:1993:GD

[Ust93c] Mikhail Ustiugov. Gold and diamonds. Bulletin of the Atomic Scientists, 49(8):51, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ustiugov:1993:IB

[Ust93d] Mikhail Ustiugov. Imaginary billions. Bulletin of the Atomic Scientists, 49(8):50, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ustiugov:1993:TNS

[Ust93e] Mikhail Ustiugov. A “temporarily nuclear state”. Bulletin of the Atomic Scientists, 49(8):33–36, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ustiugov:1996:KPP

[Ust96] Mikhail Ustiugov. Kazakh power play. Bulletin of the Atomic Scientists, 52(4):45–49, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Utley:1951:ASW

[Utl51] Clifton M. Utley. Atomic superiority — a wasting asset. Bulletin of the Atomic Scientists, 7(3):75–76, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Utroska:1987:HUN

[Utr87] Daniel Utroska. Holes in the U.S. nuclear safety net. Bulletin of the Atomic Scientists, 43(6):36–40, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonAbele:1973:CUG

[vA73] Rudolph von Abele. Communications: Unlimited growth — a good? Bulletin of the Atomic Scientists, 29(10):3, December 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Voronova-Abrams:2011:BET

[VA11] Marina Voronova-Abrams. Biosecurity 2.0: Enduring threats in the former Soviet Union. Bulletin of the Atomic Scientists, 67 (4):78–90, July/August 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/67/4/78.

Vaddi:2021:HBC

[Vad21] Pranay Vaddi. How Biden can advance nuclear arms control and stability with Russia and China. Bulletin of the Atomic Scientists, 77(1):18–20, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1859862.

Vallarta:1952:TGN

[Val52] Manual Sandoval Vallarta. The treatment of good neighbors. Bulletin of the Atomic Scientists, 8(7):250, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Vallee:1973:BBU

[Val73] Jacques Vallee. Books: The UFO Experience, by J. Allen Hynek. Bulletin of the Atomic Scientists, 29(4):49–52, April 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Valkenier:1986:BRB

[Val86] Elizabeth Kridl Valkenier. Book review: The Struggle for the Third World: Soviet Debates and American Opinion, by Jerry F. Hough. Bulletin of the Atomic Scientists, 42(8):51–52, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Valencia:1997:AD

[Val97a] Mark J. Valencia. Armed and dangerous. Bulletin of the Atomic Scientists, 53(1):52, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Valencia:1997:TW

[Val97b] Mark J. Valencia. Troubled waters. Bulletin of the Atomic Scientists, 53(1):49–54, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VanKleffens:1946:AEC

[Van46] ?. Van Kleffens. UN Atomic Energy Commission: Disarmament pacts of the past. Bulletin of the Atomic Scientists, 2(3–4):31, August 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vandenberg:1952:SVA

[Van52] Arthur H. Vandenberg. Senator Vandenberg on atomic energy. Bulletin of the Atomic Scientists, 8(5):151–156, June 1952. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VanBuitenen:1960:STF

[Van60] J. A. B. Van Buitenen. A Sanskrit tale: Four who made a lion. Bulletin of the Atomic Scientists, 16(8):310, October 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VanAllen:1961:ENS

[Van61] J. A. Van Allen. The Earth and near space. Bulletin of the Atomic Scientists, 17(5–6):218–222, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vance:1980:CUM

[Van80] Cyrus Vance. Commentary: Unmasking myths. Bulletin of the Atomic Scientists, 36(7):5–6, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vandercook:1986:SSH

[Van86] William F. Vandercook. SDI show hits the road. Bulletin of the Atomic Scientists, 42(8):16–18, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bal86b].

VanEvera:1987:BRB

[Van87] Stephen Van Evera. Book review: The Rise of the Counter-Establishment, by Sidney Blumenthal. Bulletin of the Atomic Scientists, 43(3):45–47, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vandercook:1988:BRB

[Van88] William F. Vandercook. Book review: Fighting to a Finish: The Politics of War Termination in the United States and Japan, 1945, by Leon V. Sigal. Bulletin of the Atomic Scientists, 44 (9):43–44, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VanCleave:1989:ACC

[Van89] William R. Van Cleave. Arms control criteria: Honored in the breach. Bulletin of the Atomic Scientists, 45(4):33, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VanDoren:1990:LIR

[Van90a] Charles N. Van Doren. Letter: IAEA’s role not adequately stressed. Bulletin of the Atomic Scientists, 46(8):45, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano90-59].

VanNess:1990:BRB

[Van90b] Feter Van Ness. Book review: Big bang economics: The Road to a Free Economy — Shifting from a Socialist System: The Example of Hungary, by János Kornai. Bulletin of the Atomic Scientists, 46(9):42–43, November 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VanEvera:1991:BRW

[Van91] Stephen Van Evera. Book review: When nations break up: Warpaths: the Politics of Partition, by Robert Schaeffer. Bulletin of the Atomic Scientists, 47(3):42–44, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sch91].

Vanderbilt:2003:YTR

[Van03] Tom Vanderbilt. Not your typical road trip: Book review: The Traveler’s Guide to Nuclear Weapons: A Journey Through America’s Cold War Battlefields, by James M. Maroncelli and Timothy L. Karpin. Bulletin of the Atomic Scientists, 59(6):74–75, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/6/74;   http://
www.highbeam.com/doc/1G1-110621498.html. See [MK02a].

VanDervort:2020:SLR

[Van20] Joan VanDervort. Sea level rise and beyond: Is the US military prepared for climate change? Bulletin of the Atomic Scientists, 76(3):145–149, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1751971.

Vanaik:2022:GRC

[Van22] Achin Vanaik. Global and regional confrontation in South and Southeast Asia. Bulletin of the Atomic Scientists, 78(2):78–83, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2022.2038886.

Vardanian:1996:AC

[Var96a] Astghik Vardanian. Armenia’s choice. Bulletin of the Atomic Scientists, 52(4):50–54, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Photos by Jerry Berndt.

Vardanian:1996:ALH

[Var96b] Astghik Vardanian. Armenia’s leftover heroes. Bulletin of the Atomic Scientists, 52(3):46, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vayrynen:1987:IVM

[Väy87] Raimo Väyrynen. International views: Minimum deterrence, mutual security. Bulletin of the Atomic Scientists, 43(2):45–47, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vanBeusekom:1967:CFW

[vB67] Ir. G. Th. van Beusekom. Comments: Free will and determinism. Bulletin of the Atomic Scientists, 23(8):43–45, October 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spe66].

vonBraun:1969:SAT

[vB69] Wernher von Braun. Saturn/Apollo as a transportation system. Bulletin of the Atomic Scientists, 25(7):74–78, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonBraun:1962:AF

[vBO62] Wernher von Braun and Frederick I. Ordway, III. Astronautical fallout. Bulletin of the Atomic Scientists, 18(9):13–17, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moon:1989:CW

[vCM89a] John Ellis van Courtland Moon. Chemical warfare. Bulletin of the Atomic Scientists, 45(6):41, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moon:1989:CWF

[vCM89b] John Ellis van Courtland Moon. Chemical warfare: a forgotten lesson. Bulletin of the Atomic Scientists, 45(6):40–43, July/ August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Moon:1999:DA

[vCM99a] John Ellis van Courtland Moon. Dubious allegations. Bulletin of the Atomic Scientists, 55(3):70–72, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/3/70.

Moon:1999:R

[vCM99b] John Ellis van Courtland Moon. Reviews. Bulletin of the Atomic Scientists, 55(3):70, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vanDam:1978:GPW

[vD78] André van Dam. Growth without pain: the world can achieve sustained economic growth by reducing waste. Bulletin of the Atomic Scientists, 34(4):28–30, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vandenBosch:1965:IPC

[vdB65] Robert van den Bosch. Integrated pest control in California. Bulletin of the Atomic Scientists, 21(3):22–27, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vandeHulst:1961:ISC

[vdH61] H. C. van de Hulst. International space cooperation. Bulletin of the Atomic Scientists, 17(5–6):233–236, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vegar:1997:RWT

[Veg97] Jose Vegar. Report: Weapons trade: Stiffed arms merchant sues. Bulletin of the Atomic Scientists, 53(6):12–13, November/ December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vegar:1998:CP

[Veg98a] Jose Vegar. A CBW primer. Bulletin of the Atomic Scientists, 54(2):54, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vegar:1998:TNB

[Veg98b] Jose Vegar. Terrorism’s new breed. Bulletin of the Atomic Scientists, 54(2):50–55, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vegar:1999:WS

[Veg99] Jose Vegar. Working in the shadows. Bulletin of the Atomic Scientists, 55(1):37–40, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/37.

Velikhov:1984:SWE

[Vel84] Yevgeny P. Velikhov. Space weapons: Effect on strategic stability. Bulletin of the Atomic Scientists, 40(5):[various between 32–33], May 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [OGP84].

Ventura:1979:CJT

[Ven79] Arnoldo K. Ventura. Commentary from Jamaica: Third World science and technological institutions. Bulletin of the Atomic Scientists, 35(4):49–50, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ventura:1980:BRB

[Ven80] Arnoldo K. Ventura. Book review: Progress for a Small Planet, by Barbara Ward. Bulletin of the Atomic Scientists, 36(6):50–52, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Venter:1999:ANM

[Ven99a] Al Venter. Angola: New mines, what ban? Bulletin of the Atomic Scientists, 55(3):13–15, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/3/13.

Venter:1999:NMW

[Ven99b] Al Venter. New mines, what ban? Bulletin of the Atomic Scientists, 55(3):13, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vergara:1989:LAA

[Ver89] Conzalo I. Vergara. Letter: ALCMs and ACMs. Bulletin of the Atomic Scientists, 45(6):51, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark89d].

Vest:2005:WI

[Ves05] Jason Vest. Willful ignorance. Bulletin of the Atomic Scientists, 61(4):40–48, July/August 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/4/40.

Vavilov:1948:OLD

[VFIS48] Sergei Vavilov, A. N. Frumkin, A. F. Ioffe, and N. N. Semyonov. Open letter to Dr. Einstein — from four Soviet scientists. Bulletin of the Atomic Scientists, 4(2):34, 37–38, February 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Ein48c]. Reprinted in [VFIS63].

Vavilov:1963:AES

[VFIS63] Sergei Vavilov, A. N. Frumkin, A. F. Ioffe, and N. N. Semyonov. On Albert Einstein’s support of world government. In Grodzins and Rabinowitch [GR63], pages 125–129. LCCN D842 .B78. See reply [Ein63]. Reprint of [VFIS48].

vonHippel:1958:AS

[vH58] Arthur R. von Hippel. Answers to Sputnik? Bulletin of the Atomic Scientists, 14(3):115–117, March 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1972:CC

[vH72] Frank von Hippel. Communications: A correction. Bulletin of the Atomic Scientists, 28(6):6, June 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [PvH72].

vonHippel:1975:END

[vH75a] Frank von Hippel. Editorial: The nuclear debate. Bulletin of the Atomic Scientists, 31(5):6–7, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1975:PD

[vH75b] Frank von Hippel. A perspective on the debate. Bulletin of the Atomic Scientists, 31(7):37–39, 41, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1976:CNR

[vH76a] Frank von Hippel. Correction: Nuclear reactor safety. Bulletin of the Atomic Scientists, 32(3):5, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH76b].

vonHippel:1976:NRS

[vH76b] Frank von Hippel. Nuclear reactor safety: further points of clarification. Bulletin of the Atomic Scientists, 32(1):44–45, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [vH76a].

vonHippel:1977:BJC

[vH77a] Frank von Hippel. Books: The Japanese canary: Island of Dreams: Environmental Crisis in Japan, by Norie Huddle and Michael Reich, with Nahum Stiskin. Bulletin of the Atomic Scientists, 33(1):63–64, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1977:LBR

[vH77b] Frank von Hippel. Looking back on the Rasmussen report. Bulletin of the Atomic Scientists, 33(2):42–47, February 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [SvH77, Hur77b].

vonHippel:1979:PWO

[vH79] Frank von Hippel. Perspective: Will our problems be solved if we abolish nuclear plants? Bulletin of the Atomic Scientists, 35(7):

5–6, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1980:NTP

[vH80] Frank von Hippel. The NRC and thyroid protection — one excuse after another. Bulletin of the Atomic Scientists, 36(8):44–46, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1983:METa

[vH83a] Frank von Hippel. The myths of Edward Teller. Bulletin of the Atomic Scientists, 39(3):6–12, March 1983. CODEN BASIAP. ISSN 0096-3402. See [Tel82], which is reprinted in this article, response [TvH83], and comments [Bul83, vH83b].

vonHippel:1983:METb

[vH83b] Frank von Hippel. On ‘The Myths of Edward Teller’. Bulletin of the Atomic Scientists, 39(9):60–61, November 1983. CODEN BASIAP. ISSN 0096-3402. See [vH83a].

vonHippel:1985:PAS

[vH85] Frank von Hippel. Perspective: Attacks on Star Wars critics a diversion. Bulletin of the Atomic Scientists, 41(4):8–10, April 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1987:USA

[vH87] Frank von Hippel. A U.S. scientist addresses Gorbachev. Bulletin of the Atomic Scientists, 43(4):12–13, May 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1989:PTA

[vH89] Frank von Hippel. Perspective: Taking apart the Doomsday machine. Bulletin of the Atomic Scientists, 45(4):10–12, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1990:BSD

[vH90a] Frank von Hippel. Bulletin: Soviets debate minimum deterrence. Bulletin of the Atomic Scientists, 46(1):5–6, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ray90].

vonHippel:1990:HAA

[vH^{+}90b] Frank von Hippel et al. How to avoid accidental nuclear war. Bulletin of the Atomic Scientists, 46(5):35–37, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1991:NWS

[vH91] Frank von Hippel. Nuclear weapon safety: Test ban debate, round three: Warhead safety. Bulletin of the Atomic Scientists, 47(3): 29–31, April 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [DvH91].

vonHippel:1993:RCR

[vH93] Frank von Hippel. Report: Censorship: Russian whistleblower faces jail. Bulletin of the Atomic Scientists, 49(2):7–8, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1997:A

[vH97a] Frank von Hippel. De-alerting. Bulletin of the Atomic Scientists, 53(3):35, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1997:PA

[vH97b] Frank von Hippel. Paring down the arsenal. Bulletin of the Atomic Scientists, 53(3):33–40, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Don97].

vonHippel:2000:HNC

[vH00] Frank von Hippel. Help the nuclear cities, help ourselves. Bulletin of the Atomic Scientists, 56(6):21–23, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/6/21.

vonHippel:2001:WFS

[vH01] Frank N. von Hippel. Where Fermi stood. Bulletin of the Atomic Scientists, 57(5):26–29, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/5/26;
http://bos.sagepub.com/content/57/5/26.full;   http:/
/www.highbeam.com/doc/1G1-78334537.html.

vonHippel:2011:RPC

[vH11] Frank N. von Hippel. The radiological and psychological consequences of the Fukushima Daiichi accident. Bulletin of the Atomic Scientists, 67(5):27–36, September/October 2011. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/67/5/27.

vonHippel:2016:HMN

[vH16] Frank N. von Hippel. How many nuclear warheads does the United States need? Bulletin of the Atomic Scientists, 72(5): 274–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1976:LRE

[vHC76] Frank von Hippel and David Dinsmore Comey. Letter: Rapid evacuation. Bulletin of the Atomic Scientists, 32(3):2–3, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com75b].

vonHippel:1986:CES

[vHC86] Frank von Hippel and Thomas B. Cochran. Chernobyl: the emerging story: Estimating long-term health effects. Bulletin of the Atomic Scientists, 42(7):18–24, August/September 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [WvH86].

vonHippel:1996:THL

[vHJ96] Frank von Hippel and Suzanne Jones. Take a hard look at subcritical tests. Bulletin of the Atomic Scientists, 52(6):44–47, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Gar97b].

vonHippel:1997:BB

[vHJ97a] Frank von Hippel and Suzanne Jones. Birth of the breeder. Bulletin of the Atomic Scientists, 53(5):48–49, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1997:SDF

[vHJ97b] Frank von Hippel and Suzanne Jones. The slow death of the fast breeder. Bulletin of the Atomic Scientists, 53(5):46–51, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1975:ST

[vHW75] Frank von Hippel and Robert H. Williams. Solar technologies. Bulletin of the Atomic Scientists, 31(9):25–31, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [vHWF^{+}81].

vonHippel:1976:EWN

[vHW76a] Frank von Hippel and Robert H. Williams. Energy waste and nuclear power growth. Bulletin of the Atomic Scientists, 32(10): 18–21, 48–56, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1976:EWC

[vHW76b] Frank von Hippel and Robert H. Williams. Energy waste in the COAT scenario. Bulletin of the Atomic Scientists, 32(10):49, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1977:TSC

[vHW77] Frank von Hippel and Robert H. Williams. Toward a solar civilization. Bulletin of the Atomic Scientists, 33(8):12–15, 56–60, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonHippel:1976:LTC

[vHWD76] Frank von Hippel, Robert H. Williams, and H. C. Dudley. Letter: Taxes credulity. Bulletin of the Atomic Scientists, 32(1):2–3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75b] and comment [Nov76].

vonHippel:1981:CNQ

[vHWF^{+}81] Frank von Hippel, Robert H. Williams, Bernard T. Feld, Englebert Broda, and Bent Sørensen. Commentary: ‘No quick fix’. Bulletin of the Atomic Scientists, 37(6):63–64, June/July 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vHW75, KB80].

vonHippel:1979:TCP

[vHWL79] Frank von Hippel, Robert H. Williams, and Amory B. Lovins. On thorium cycles and proliferation. Bulletin of the Atomic Scientists, 35(5):50–54, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lov79].

vonHippel:1993:NJT

[vHZC93] Frank von Hippel and Tom Zamora-Collina. Nuclear junkies: Testing, testing, 1, 2, 3 — forever. Bulletin of the Atomic Scientists, 49(6):28–32, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vickers:1978:FDR

[Vic78] T. Vickers. Flexible DNA regulation: the British model. Bulletin of the Atomic Scientists, 34(1):4, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vid:1990:GBS

[Vid90] Leonid Vid. Guns into butter, Soviet style. Bulletin of the Atomic Scientists, 46(1):16–19, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Viederman:1962:AEN

[Vie62] Stephen Viederman. Academic exchange — a narrow bridge. Bulletin of the Atomic Scientists, 18(2):17–21, February 1962. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Villa:1975:CSJ

[Vil75] Brian Loring Villa. Confusion of signals: James Franck, the Chicago scientists, and early efforts to stop the bomb. Bulletin of the Atomic Scientists, 31(10):36–43, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vinogradov:1959:PPM

[Vin59] A. P. Vinogradov. Prospects for the Pugwash Movement. Bulletin of the Atomic Scientists, 15(9):376–378, November 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Visscher:1956:PVS

[Vis56] Maurice B. Visscher. Part VII: Science and religion: The duty to doubt and the will to believe. Bulletin of the Atomic Scientists, 12(10):356–359, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Visscher:1959:LER

[Vis59] Maurice B. Visscher. Letter to the Editor: Religion without revelation? Bulletin of the Atomic Scientists, 15(10):433, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor59a].

Jehle:1969:NKJ

[vKJ69] Dietlinde von Kuenssberg Jehle. Note on Karl Jaspers. Bulletin of the Atomic Scientists, 25(3):36, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonLaue:1947:KDP

[vL47] Max von Laue. Die Kriegstätigkeit der deutschen Physiker. (German) [The war work of the German physicist]. Physikalische Blätter, 3(12):424–425, December 1947. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/phbl.19470031206/

abstract. Response to [Mor47a]. See also reply [Mor48a].

vonLaue:1948:KDP

[vL48a] Max von Laue. Die Kriegsaktivitätkeit der deutschen Physiker. (German) [The wartime activities of the German physicist]. Physikalische Blätter, 4(??):424–425, ???? 1948. CODEN PH-BLAG. ISSN 0031-9279 (print), 1521-3722 (electronic). Cited vaguely in [Wal90b, page 59].

vonLaue:1948:FDR

[vL48b] Max von Laue. Freedom demands responsibility — the scientist’s role in society: The world needs new ethical standards. Bulletin of the Atomic Scientists, 4(11):337–338, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a].

vonLaue:1948:WAG

[vL48c] Max von Laue. The wartime activities of German scientists. Bulletin of the Atomic Scientists, 4(4):103, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See reply [Mor48a], comment [Rab48d], and further reply [vL48a].

vonLaue:1963:MSO

[vL63] Theodore H. von Laue. Modern science and the Old Adam. Bulletin of the Atomic Scientists, 19(1):2–5, January 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and replies [JvL63, GC63, Lar63].

vonLaue:1965:FSW

[vL65] Theodore von Laue. Forum: The subversive West. Bulletin of the Atomic Scientists, 21(5):25–28, May 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonLaue:1985:LSE

[vL85] Theodore H. von Laue. Letter: Sakharov and East–West relations. Bulletin of the Atomic Scientists, 41(1):54, January 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano84e, BP85].

vonLaue:1986:LHR

[vL86] Theodore H. von Laue. Letter: Human rights in the Soviet Union. Bulletin of the Atomic Scientists, 42(6):63–64, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gra86].

vonLohmann:1992:BRN

[vL92] Fred von Lohmann. Book review: National secrecy council: Keepers of the Keys: A History of the NSC from Truman to Bush, by John Prados. Bulletin of the Atomic Scientists, 48(2):43–44, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonMuller:1989:ACC

[vM89] Albrecht A. C. von Müller. Arms control criteria: More to do. Bulletin of the Atomic Scientists, 45(4):43–44, May 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vanMoyland:1998:PT

[vMC98] Suzanna van Moyland and Roger Clark. The paper trail: Seismographs all over the world jumped on May 11, but not on May 13. Bulletin of the Atomic Scientists, 54(4):26–29, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonMuller:1988:EWN

[vMK88] Albrecht A. C. von Müller and Andrzej Karkoszka. An East– West negotiating proposal. Bulletin of the Atomic Scientists, 44 (7):39–41, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [KWM88].

vonNeumann:1957:TBC

[vN57] John von Neumann. Testimony before the Congressional Committee on Atomic Energy, January 31, 1946. Bulletin of the Atomic Scientists, 13(4):151, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vofsi:1972:AIC

[Vof72] David Vofsi. An agro-industrial complex in the Arava: Bringing an arid land to life. Bulletin of the Atomic Scientists, 28(8):45–51, October 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on science in a small developing country — Israel.

Voge:1987:LBP

[Vog87] Hervey H. Voge. Letter: Bulletin policy. Bulletin of the Atomic Scientists, 43(2):59–60, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [AG86, Dum86, Cro86].

Vogel:1990:JSI

[Vog90a] Steven K. Vogel. Just say “iee” [no]. Bulletin of the Atomic Scientists, 46(1):34, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vogel:1990:NWL

[Vog90b] Steven K. Vogel. New weapon label: Made in Japan. Bulletin of the Atomic Scientists, 46(1):31–35, January/February 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vogel:2013:IAP

[Vog13] Kathleen M. Vogel. Intelligent assessment: Putting emerging biotechnology threats in context. Bulletin of the Atomic Scientists, 69(1):43–52, January/February 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/69/1/43.

Vohra:1970:INP

[Voh70] Hans R. Vohra. India’s nuclear policy of three negatives. Bulletin of the Atomic Scientists, 26(4):25–27, April 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vohra:1971:EPA

[Voh71a] Hans R. Vohra. Environmental Protection Agency: Teething troubles. Bulletin of the Atomic Scientists, 27(4):30–31, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vohra:1971:LWM

[Voh71b] Hans R. Vohra. A letter from Washington: Ministry of Science — U.S. style. Bulletin of the Atomic Scientists, 27(1):29–32, January 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vohra:1971:LWP

[Voh71c] Hans R. Vohra. Letter from Washington: Preventive technology: Rival proposals in Congress. Bulletin of the Atomic Scientists, 27(3):25–26, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vohra:1971:LWS

[Voh71d] Hans R. Vohra. Letter from Washington: Science in Congress. Bulletin of the Atomic Scientists, 27(2):44–56, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Volborth:1974:OMC

[Vol74] Alex Volborth. Oxygen in the Moon’s crust: A glimpse of present lunar research from a narrow perspective. Bulletin of the Atomic Scientists, 30(3):35–41, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

VonLaue:1983:HRI

[Von83] Theodore H. Von Laue. Human rights imperialism. Bulletin of the Atomic Scientists, 39(7):32–35, August/September 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mac83, EV84, Fri84].

Vosburg:1988:LGE

[Vos88] Albert C. Vosburg. Letter: GITSM every time. Bulletin of the Atomic Scientists, 44(8):52–53, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Voyles:1976:LDB

[Voy76] D. E. Voyles. Letter: Delicate balance. Bulletin of the Atomic Scientists, 32(2):3, February 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Viorst:1967:RDF

[VR67] Milton Viorst and J. V. Reistrup. Radon daughters and the Federal government. Bulletin of the Atomic Scientists, 23(8):25–29, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonSuttner:1978:PNP

[vS78] Bertha von Suttner. Perspective: Nobel Peace Laureates: 1905. Bulletin of the Atomic Scientists, 34(8):1, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From Nobel Peace Prize Lecture, 18 April 1906, Oslo, Norway.

Vilensky:2004:CR

[VS04a] Joel A. Vilensky and Pandy R. Sinish. Cleanup in Russia. Bulletin of the Atomic Scientists, 60(2):58–59, March/April 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vilensky:2004:DD

[VS04b] Joel A. Vilensky and Pandy R. Sinish. The dew of death. Bulletin of the Atomic Scientists, 60(2):54–60, March/April 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/60/2/54.

Vorosmarty:2009:BSW

[VSD^{+}09] Charles J. Vörösmarty, James Syvitski, John Day, Alex de Sherbinin, Liviu Giosan, and Chris Paola. Battling to save the World’s river deltas. Bulletin of the Atomic Scientists, 65(2): 31–43, March/April 2009. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/65/2/31.

Vallentyne:1973:DGW

[VT73] J. R. Vallentyne and H. L. Tracy. Demophora: Greeks had words for the interdependence of biological and technical growth. Bulletin of the Atomic Scientists, 29(5):24, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Vestergaard:2022:BBC

[VU22] Cindy Vestergaard and Lovely Umayam. Blockchain beyond cryptocurrency: A revolution in information management and international security. Bulletin of the Atomic Scientists, 78(4): 198–202, 2022. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2022.2090116.

Vucinich:1969:BBI

[Vuc69] Alexander Vucinich. Books: Science and Ideology in Soviet Society, edited by George Fischer. Bulletin of the Atomic Scientists, 25(3):33, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vanValkengoed:2023:CAM

[vV23] Anne M. van Valkengoed. Climate anxiety is not a mental health problem. but we should still treat it as one. Bulletin of the Atomic Scientists, 79(6):385–387, 2023. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2023.2266942.

vonWeizsacker:1957:SGA

[vW57] Carl Friedrich von Weizsäcker. Should Germany have atomic arms. Bulletin of the Atomic Scientists, 13(8):283–286, October 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Followup on [BBF^{+}57].

vonWeizsacker:1958:DWW

[vW58] Carl Friedrich von Weizsäcker. Do we want to save ourselves? Bulletin of the Atomic Scientists, 14(5):180–184, May 1958. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonWeizsacker:1959:DG

[vW59] Carl Friedrich von Weizsäcker. Disengagement in Germany? Bulletin of the Atomic Scientists, 15(6):226–228, June 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

vonWeizsacker:1980:EA

[vW80] C. F. von Weizsäcker. European armaments in the 1980s. Bulletin of the Atomic Scientists, 36(10):8–11, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Translation by Bulletin staff of German original published in the 16 November 1979 issue of Die Zeit.

Winslow:1993:LEE

[WA93] Robert E. Winslow and William M. Arkin. Letter: Not even the experts know …. Bulletin of the Atomic Scientists, 49(6):55, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark93b].

Waddington:1958:STC

[Wad58] C. H. Waddington. Scientific and technical cooperation: A key to mutual trust. Bulletin of the Atomic Scientists, 14(6):212–214, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waddington:1963:ICM

[Wad63] C. H. Waddington. International cooperation: Mobilizing the world’s biologists. Bulletin of the Atomic Scientists, 19(9):39– 41, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wadia:1974:IPS

[Wad74] Avabai B. Wadia. India’s painful struggle toward the ‘small family norm’: One billion mouths to feed by the turn of the century. Bulletin of the Atomic Scientists, 30(6):25–28, June 1974. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wagner:1961:BBS

[Wag61] Philip L. Wagner. Books: After the Seventh Day: the World Man Created, by Ritchie Calder. Bulletin of the Atomic Scientists, 17 (10):436–437, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wagner:1975:TM

[Wag75] Thomas P. Wagner. Trivial minutiae. Bulletin of the Atomic Scientists, 31(6):4, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom75a].

Wagoner:1983:D

[Wag83] David Wagoner. In distress. Bulletin of the Atomic Scientists, 39(1):4–5, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Poem constructed from the International Code of Signals.

Wagner:1988:MWC

[Wag88] Richard L. Wagner, Jr. Making warheads: A case for producing nuclear material. Bulletin of the Atomic Scientists, 44(1):43– 45, January/February 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wain:1960:SAM

[Wai60] John Wain. A song about Major Eatherly. Bulletin of the Atomic Scientists, 16(9):360–361, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). First published in The Listener on 6 August 1959. Major Claude R. Eatherly piloted the plane that dropped the second atomic bomb on Japan, on Nagasaki on 9 August 1945. See [Gig60].

Wallace:1946:LP

[Wal46a] Henry A. Wallace. From the letter to the President. Bulletin of the Atomic Scientists, 2(7–8):2–3, October 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wallace:1946:HWS

[Wal46b] Henry A. Wallace. Henry A. Wallace supports the McMahon Bill. Bulletin of the Atomic Scientists of Chicago, 1(5):6–7, February 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wallace:1950:UUF

[Wal50] Henry A. Wallace. The U.S., the U.N., and Far Eastern agriculture. Bulletin of the Atomic Scientists, 6(12):363–364, December 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1951:EGS

[Wal51] Eric A. Walker. Effects of government support on scientific research. Bulletin of the Atomic Scientists, 7(4):119–121, April 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wallace:1952:MLS

[Wal52] Henry A. Wallace. The Maginot line of secrecy. Bulletin of the Atomic Scientists, 8(2):36–37, February 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walkowicz:1955:LEA

[Wal55] T. F. Walkowicz. Letter to the Editor: Air Force Association convention. Bulletin of the Atomic Scientists, 11(5):201, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walford:1961:HS

[Wal61] Lionel Walford. Harvest from the sea. Bulletin of the Atomic Scientists, 17(10):415–418, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1964:ASR

[Wal64] George Richmond Walker. Art, science, and reality. Bulletin of the Atomic Scientists, 20(7):9–12, September 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wald:1969:CSC

[Wal69] George Wald. Chicago: The social control of science and its applications: America’s my home, not my business, my home. Bulletin of the Atomic Scientists, 25(5):29–31, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wald:1974:EAY

[Wal74] George Wald. Editorial: Arise, ye pioneers. Bulletin of the Atomic Scientists, 30(10):4–6, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mat75].

Walters:1976:LPV

[Wal76] Robert E. Walters. Letter: Policy vs. strategy. Bulletin of the Atomic Scientists, 32(1):3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel75a].

Wales:1977:LCV

[Wal77] Jack V. Wales, Jr. Letter: Certainty vs. truth. Bulletin of the Atomic Scientists, 33(2):4, 6, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei76b, Bru76].

Walker:1981:BRB

[Wal81a] Paul F. Walker. Book review: MX: Prescription for Disaster, by Herbert Scoville, Jr.; Misguided eXpenditure, by David Gold, Christopher Paine and Gail Shields. Bulletin of the Atomic Scientists, 37(10):34–36, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walters:1981:CPN

[Wal81b] Robert E. Walters. Commentary: On ‘Preventing Nuclear War’. Bulletin of the Atomic Scientists, 37(8):64–65, October 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fis81].

Walker:1982:TIA

[Wal82] Paul F. Walker. Teach-ins on American campuses. Bulletin of the Atomic Scientists, 38(2):10–11, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walsh:1983:TMI

[Wal83a] Edward J. Walsh. Three Mile Island: meltdown of democracy? Bulletin of the Atomic Scientists, 39(3):57–60, March 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BW83].

Walsh:1983:BRB

[Wal83b] Edwin J. Walsh. Book review: The Atom Besieged: Antinuclear Movements in France and Germany, by Dorothy Nelkin and Michael Pollak. Bulletin of the Atomic Scientists, 39(7):50, August/September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walsh:1984:BRB

[Wal84] Edward J. Walsh. Book review: The Next Nuclear Gamble: Transportation and Storage of Nuclear Waste, by Marvin Resnikoff. Bulletin of the Atomic Scientists, 40(3):45, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1985:BRB

[Wal85a] Paul F. Walker. Book review: Hawks, Doves, and Owls: An Agenda for Avoiding Nuclear War, edited by Graham T. Allison, Albert Carnesale, and Joseph S. Nye, Jr. Bulletin of the Atomic Scientists, 41(11):59–60, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walsh:1985:TMI

[Wal85b] Edward J. Walsh. Three Mile island: the battle of Unit 1. Bulletin of the Atomic Scientists, 41(5):30–31, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walworth:1986:LCL

[Wal86] Edward H. Walworth, Jr. Letter: A Challenger lesson. Bulletin of the Atomic Scientists, 42(4):56, April 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mye86].

Walker:1987:BRB

[Wal87a] Paul F. Walker. Book review: The Price of Peace, by Lawrence Freedman. Bulletin of the Atomic Scientists, 43(1):56, January/ February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walt:1987:WRP

[Wal87b] Stephen M. Walt. Washington report: Politicians and professors: a double standard. Bulletin of the Atomic Scientists, 43(6):3–4, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [War87].

Walker:1989:ACC

[Wal89a] Paul F. Walker. Arms control criteria: Limited success. Bulletin of the Atomic Scientists, 45(4):40, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1989:LDA

[Wal89b] Paul F. Walker. Letter: David ACDA? Bulletin of the Atomic Scientists, 45(4):52, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kre88b].

Walker:1989:SUD

[Wal89c] Paul F. Walker. START update: Did anyone tell the Pentagon? Bulletin of the Atomic Scientists, 45(8):12, 40, October 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1990:DUB

[Wal90a] J. Samuel Walker. The decision to use the bomb: A historiographical update. Diplomatic History, 14(1):97–114, January 1990. CODEN ???? ISSN 0145-2096 (print), 1467-7709 (electronic).

Walker:1990:HGG

[Wal90b] Mark Walker. Heisenberg, Goudsmit and the German Atomic Bomb. Physics Today, 43(1):52–60, January/February 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/

phtoad/v43/i1/p52_s1. See responses and rebuttal [LRD^{+}91].

Walker:1990:FPE

[Wal90c] Paul F. Walker. France’s popular Exocet. Bulletin of the Atomic Scientists, 46(4):24, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1990:HTK

[Wal90d] Paul F. Walker. High-tech killing power. Bulletin of the Atomic Scientists, 46(4):23–26, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walsh:1990:BBW

[Wal90e] Jacquelyn Walsh. Bulletin: Big weapon in the antidrug war. Bulletin of the Atomic Scientists, 46(5):5–6, June 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walt:1999:MLU

[Wal99] Stephen M. Walt. Musclebound: The limits of U.S. power. Bulletin of the Atomic Scientists, 55(2):44–48, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/2/44.

Walker:2006:CS

[Wal06] William Walker. Cut off the source. Bulletin of the Atomic Scientists, 62(2):68, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/2/68.

Wald:2007:GPP

[Wal07] Matthew L. Wald. Getting power to the people. Bulletin of the Atomic Scientists, 63(5):??, September 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/5/26.

Wallace:2008:NWF

[Wal08] Michael D. Wallace. A nuclear-weapon-free Arctic. Bulletin of the Atomic Scientists, 64(1):60, March/April 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/1/60.

Walker:2021:HNP

[Wal21a] William Walker. The history of nuclear power’s imagined future: Plutonium’s journey from asset to waste. Bulletin of the Atomic Scientists, 77(5):259–264, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2021.
1964257.

Walsh:2021:JTU

[Wal21b] Jason Walsh. A just transition for US workers is within reach. Bulletin of the Atomic Scientists, 77(2):86–89, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1886447.

Wang:1983:CNP

[Wan83] Chi Wang. China’s nuclear programs and policies. Bulletin of the Atomic Scientists, 39(3):18–21, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

WH:1996:CNT

[Wan96] Wang Hao. China is no threat. Bulletin of the Atomic Scientists, 52(4):19–21, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wanderer:2005:TW

[Wan05] William R. Wanderer, Jr. The tools but not the will. Bulletin of the Atomic Scientists, 61(1):70–71, January/February 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/1/70.

Wang:2016:PPM

[Wan16] Haibin Wang. Population’s part in mitigating climate change: A Chinese response. Bulletin of the Atomic Scientists, 72(3):192– 193, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warren:1948:MPA

[War48] Shields Warren. The medical program of the Atomic Energy Commission. Bulletin of the Atomic Scientists, 4(8):233–234, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warren:1949:BBI

[War49] Shields Warren. Books: Histopathology of Irradiation from External and Internal Sources, William Bloom. Bulletin of the Atomic Scientists, 5(6–7):204, June/July 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warne:1959:LEA

[War59] Clore Warne. Letter to the Editor: After missiles and satellites, what? Bulletin of the Atomic Scientists, 15(5):225, May 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab57b, Pri58b].

Warburg:1963:TCD

[War63] James P. Warburg. Today’s challenges to diplomacy. Bulletin of the Atomic Scientists, 19(9):2–4, November 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ward:1968:TCW

[War68] Barbara Ward. Technological change and the world market. Bulletin of the Atomic Scientists, 24(6):8–12, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warnke:1978:ACG

[War78] Paul C. Warnke. Arms control: a global imperative. Bulletin of the Atomic Scientists, 34(6):32–34, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warnke:1986:BRB

[War86] Paul C. Warnke. Book review: Secrets of State, by Barry Rubin. Bulletin of the Atomic Scientists, 42(1):48–50, January 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warshawsky:1987:LEA

[War87] Irving Warshawsky. Letter: Ethics in academia. Bulletin of the Atomic Scientists, 43(8):54, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wal87b].

Warnke:1988:ASI

[War88] Paul C. Warnke. After the summit: INF Treaty: a good start. Bulletin of the Atomic Scientists, 44(2):18–20, March 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warnke:1989:ACC

[War89] Paul C. Warnke. Arms control criteria: A little shoring up. Bulletin of the Atomic Scientists, 45(4):35, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warnke:1991:SSL

[War91] Paul C. Warnke. START: Success linked to ABM treaty. Bulletin of the Atomic Scientists, 47(9):18–19, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warnke:1992:BRN

[War92a] Paul C. Warnke. Book review: Nuclear Israel: The Samson Option, by Seymour M. Hersch. Bulletin of the Atomic Scientists, 48(2):41–42, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warnke:1992:MM

[War92b] Paul C. Warnke. Missionless missiles. Bulletin of the Atomic Scientists, 48(4):36–38, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Warf:1995:LVS

[War95] James C. Warf. Letter: A visit to Semipalatinsk. Bulletin of the Atomic Scientists, 51(5):60–61, September/October 1995. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wara:2014:CEC

[War14] Michael Wara. California’s energy and climate policy: a full plate, but perhaps not a model policy. Bulletin of the Atomic Scientists, 70(5):26–34, September 2014. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waskow:1962:BBD

[Was62] Arthur I. Waskow. Books: Conflict and Defense: A General Theory, by Kenneth E. Boulding. Bulletin of the Atomic Scientists, 18(7):30–31, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waskow:1963:LEC

[Was63] Arthur I. Waskow. Letter to the Editor: Contradictory initiatives. Bulletin of the Atomic Scientists, 19(4):35, April 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waskow:1964:AAN

[Was64] Arthur I. Waskow. Advancing the American national interest without war. Bulletin of the Atomic Scientists, 20(2):23–25, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment and reply [FW64].

Wasawo:1966:DAS

[Was66] David P. S. Wasawo. Developing the academic spirit in East Africa. Bulletin of the Atomic Scientists, 22(8):19–21, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Washburn:1982:FYS

[Was82] Sherwood Washburn. Fifty years of studies on human evolution. Bulletin of the Atomic Scientists, 38(5):37–43, May 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Washburn:1997:WMT

[Was97] Jennifer Washburn. When money talks, Congress listens. Bulletin of the Atomic Scientists, 53(4):38–41, July/August 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waterman:1951:PRN

[Wat51] Alan T. Waterman. Present role of the National Science Foundation. Bulletin of the Atomic Scientists, 7(6):165–167, June 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waterman:1953:RND

[Wat53] Alan T. Waterman. Research for national defense. Bulletin of the Atomic Scientists, 9(2):36–38, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waterman:1957:BRN

[Wat57] Alan T. Waterman. Basic research — a national resource. Bulletin of the Atomic Scientists, 13(10):369–371, December 1957.

CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waterman:1963:SSM

[Wat63] Alan T. Waterman. Science in the service of man. Bulletin of the Atomic Scientists, 19(5):3–6, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Watson:1970:CNW

[Wat70] William W. Watson. Communications: Nuclear weapons for U.N.? Bulletin of the Atomic Scientists, 26(3):48, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Waters:1973:LMD

[Wat73a] W. G. Waters II. Landing a man downtown. Bulletin of the Atomic Scientists, 29(9):34–35, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Watts:1973:CLM

[Wat73b] Robert G. Watts. Communications: “Limits”: Misunderstood. Bulletin of the Atomic Scientists, 29(9):2–3, November 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Har72a, Ber72b].

Wattenberg:1974:AOT

[Wat74] Albert Wattenberg. ‘All in our time’: The building of the first chain reaction pile. Bulletin of the Atomic Scientists, 30(6):51–57, June 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Watson:1977:IMO

[Wat77] James D. Watson. An imaginary monster: the only danger we face is the specter of untested regulations. Bulletin of the Atomic Scientists, 33(5):12–13, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wattenberg:1982:DEP

[Wat82] Albert Wattenberg. December 2, 1942: the event and the people. Bulletin of the Atomic Scientists, 38(10):22–32, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wattenberg:1992:LE

[Wat92] Albert Wattenberg. A lovely experiment. Bulletin of the Atomic Scientists, 48(10):41–43, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Watterson:2019:WNS

[Wat19] Christopher J. Watterson. What next for sanctions against North Korea? Bulletin of the Atomic Scientists, 75(5):247–251, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2019.1654270.

Waymack:1951:FYU

[Way51] W. W. Waymack. Four years under the law. Bulletin of the Atomic Scientists, 7(2):51–56, February 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wayman:1965:HLS

[Way65] Cooper H. Wayman. A hard look at soft detergents. Bulletin of the Atomic Scientists, 21(4):23–26, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1951:QC

[WB51] Charles E. Wilson and Hanson W. Baldwin. Quotes without comment. Bulletin of the Atomic Scientists, 7(9):288, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1985:LDP

[WB85a] Alvin M. Weinberg and Jack N. Barkenbus. Letter: Defenseprotected build-down. Bulletin of the Atomic Scientists, 41(2): 52, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westing:1985:LNH

[WB85b] Arthur H. Westing and Paul Boyer. Letter: Nuclear history. Bulletin of the Atomic Scientists, 41(1):56–57, January 1985. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Boy84].

Waller:1987:WRH

[WB87] Douglas C. Waller and James T. Bruce. Washington report: Holes in the impenetrable shield. Bulletin of the Atomic Scientists, 43(8):5–6, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weismiller:1995:LFY

[WB95] Frances P. Weismiller and John Bradley. Letters: Fifty years later. Bulletin of the Atomic Scientists, 51(5):59, September/ October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Moo95e].

Watson-Boles:2006:ECD

[WB06] Amanda Watson-Boles. Europe’s (not so) continental divide. Bulletin of the Atomic Scientists, 62(4):62–63, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/62/4/62.

Waterman:1947:RRO

[WC47] Alan T. Waterman and Robert D. Conrad. Replies to Ridenour: Office of Naval Research discusses Ridenour’s views. Bulletin of the Atomic Scientists, 3(8):230, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weeks:1973:TII

[WC73] W. F. Weeks and W. J. Campbell. Towing icebergs to irrigate arid lands: Manna or madness? Bulletin of the Atomic Scientists, 29(5):35–39, May 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lon74] and response [WC74].

Weeks:1974:CR

[WC74] W. F. Weeks and W. J. Campbell. Communications: Response. Bulletin of the Atomic Scientists, 30(3):2, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WC74, Lon74].

Weber:2017:MSB

[WC17] Steven Weber and Betsy Cooper. Moving slowly, not breaking enough: Trump’s cybersecurity accomplishments. Bulletin of the Atomic Scientists, 73(6):388–394, 2017. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2017.1388676.

Wheelis:2003:BB

[WD03] Mark Wheelis and Malcolm Dando. Back to bioweapons. Bulletin of the Atomic Scientists, 59(1):40–46, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/1/40.

Weaver:1957:SC

[Wea57] Warren Weaver. Science and the citizen. Bulletin of the Atomic Scientists, 13(10):361–365, December 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Published simultaneously in Science, 13 December 1957.

Weaver:1963:DR

[Wea63] Warren Weaver. Dreams and responsibilities. Bulletin of the Atomic Scientists, 19(5):10–11, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weatherwax:1975:VLR

[Wea75] Robert K. Weatherwax. Virtues and limitations of risk analysis. Bulletin of the Atomic Scientists, 31(7):29–37, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wear:1976:LBD

[Wea76] R. Wear. Letter: The bomb has dropped. Bulletin of the Atomic Scientists, 32(9):4, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weart:1979:SPF

[Wea79] Spencer R. Weart. Scientists in power: France and the origins of nuclear energy, 1900–1950cg. Bulletin of the Atomic Scientists, 35(3):41–50, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpts from the author’s 1979 book Scientists in Power.

Weart:1982:FRB

[Wea82] Spencer R. Weart. Film review: The Day After Trinity: J. Robert Oppenheimer and the Atomic Bomb, produced and directed by John Else. Bulletin of the Atomic Scientists, 38(7):41–42, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weart:1985:AAH

[Wea85] Spencer R. Weart. The Atomic Age: The heyday of myth and cliché. Bulletin of the Atomic Scientists, 41(7):38–43, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weart:1992:NFP

[Wea92] Spencer Weart. From the nuclear frying pan into the global fire. Bulletin of the Atomic Scientists, 48(5):18–27, June 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1999:WFC

[Wea99] Jim Williams and et al. The wind farm in the cabbage patch. Bulletin of the Atomic Scientists, 55(3):40–48, May/June 1999. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/55/3/40.

Weart:2011:GWH

[Wea11] Spencer Weart. Global warming: How skepticism became denial. Bulletin of the Atomic Scientists, 67(1):41–50, January/February 2011. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/67/1/41.

Webb:1993:CU

[Web93] Gregory Webb. CFE update. Bulletin of the Atomic Scientists, 49(10):24, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Webster:2002:MGT

[Web02] Paul Webster. Minatom: The grab for trash. Bulletin of the Atomic Scientists, 58(5):33–37, September/October 2002. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/58/5/33.

Webster:2003:JLO

[Web03] Paul Webster. Just like old times. Bulletin of the Atomic Scientists, 59(4):30–35, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/59/4/30.

Webster:2006:NPN

[Web06] Paul Webster. Nuclear power: No can-do? Bulletin of the Atomic Scientists, 62(1):18–19, January/February 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/1/18.

Weeks:1973:CGP

[Wee73] W. F. Weeks. Communications: Geophysical problems. Bulletin of the Atomic Scientists, 29(4):54–55, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, ZSA73b].

Weeks:1995:LSA

[Wee95] Jennifer Weeks. Letter: Spratt amendment was supported. Bulletin of the Atomic Scientists, 51(5):60, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa95b].

Weeks:1998:WOL

[Wee98] Jennifer Weeks. Report: Energy Department: Will O’Leary legacy last? Bulletin of the Atomic Scientists, 54(2):11–14, March/April 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weeks:2003:WW

[Wee03] Jennifer Weeks. Why not wind? Bulletin of the Atomic Scientists, 59(1):70–71, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/1/70.

Weglarczyk:1996:BJL

[Weg96a] Bartosz Weglarczyk. Bulletin: Just like the bad old days? Bulletin of the Atomic Scientists, 52(5):5–7, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weglarczyk:1996:RNS

[Weg96b] Bartosz Weglarczyk. Report: Nuclear smuggling: Smoke, but no fire. Bulletin of the Atomic Scientists, 52(5):15, September/ October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1946:LEO

[Wei46] Victor F. Weisskopf. Letter to the Editor: Our publication policy as seen from abroad. Bulletin of the Atomic Scientists, 2(11–12): 11, December 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1948:BRB

[Wei48] Victor F. Weisskopf. Book review: Die Geschichte der Atombombe [(German) The History of the Atomic Bomb], by Hans Thirring. Bulletin of the Atomic Scientists, 4(4):105, April 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1952:EVP

[Wei52a] Victor F. Weisskopf. From the Editors: Visa procedures show improvement: Role of scientific advisors. Bulletin of the Atomic Scientists, 8(8):293, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1952:RVS

[Wei52b] Victor F. Weisskopf. Report on the visa situation. Bulletin of the Atomic Scientists, 8(7):221–222, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Weil:1953:RPN

[Wei53a] George L. Weil. A role for private nonprofit organizations. Bulletin of the Atomic Scientists, 9(5):165–168, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1953:HSW

[Wei53b] Alvin M. Weinberg. How shall we establish a nuclear power industry in the United States? Bulletin of the Atomic Scientists, 9(4): 120–124, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiner:1953:PF

[Wei53c] Betsy P. Weiner. Playing with fire. Bulletin of the Atomic Scientists, 9(4):147, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1954:NJT

[Wei54a] Alvin M. Weinberg. A nuclear journey through Europe. Bulletin of the Atomic Scientists, 10(6):215–217, June 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1954:VFS

[Wei54b] Victor F. Weisskopf. Visas for foreign scientists. Bulletin of the Atomic Scientists, 10(3):68–69, 112, March 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weil:1955:APG

[Wei55a] George L. Weil. American plans for Geneva conference. Bulletin of the Atomic Scientists, 11(6):217–219, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1955:BRB

[Wei55b] V. F. Weisskopf. Book review: Two Minutes to Midnight, by Elmer Davis. Bulletin of the Atomic Scientists, 11(6):226, June 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1955:TPG

[Wei55c] Victor Weisskopf. A theoretical physicist at the Geneva Conference. Bulletin of the Atomic Scientists, 11(8):278–279, October 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1956:PIS

[Wei56a] Alvin M. Weinberg. Part IV: Science and world economy: Today’s revolution. Bulletin of the Atomic Scientists, 12(8):299–302, October 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Ano56f].

Weisskopf:1956:ICH

[Wei56b] V. F. Weisskopf. International Conference on High-Energy Physics. Bulletin of the Atomic Scientists, 12(7):258–260, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1958:TPC

[Wei58] Alvin M. Weinberg. Third Pugwash Conference: Prospects in international science. Bulletin of the Atomic Scientists, 14(10):

402–404, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1959:GST

[Wei59a] Alvin M. Weinberg. Geneva, 1958. some thoughts on reactors. Bulletin of the Atomic Scientists, 15(3):132–137, March 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:1959:MS

[Wei59b] Paul Weiss. The message of science. Bulletin of the Atomic Scientists, 15(7):274–277, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinhaus:1961:RCS

[Wei61] Robert S. Weinhaus. Readers’ comments: A self-enforcing treaty. Bulletin of the Atomic Scientists, 17(4):168–169, April 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weidenbaum:1964:OC

[Wei64a] Murray L. Weidenbaum. Obstacles to conversion. Bulletin of the Atomic Scientists, 20(4):10–14, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1964:FNE

[Wei64b] Alvin M. Weinberg. Forum: The new estate. Bulletin of the Atomic Scientists, 20(2):16–19, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1965:WPS

[Wei65] Victor F. Weisskopf. Why pure science? Bulletin of the Atomic Scientists, 21(4):4–8, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [RB65].

Weinberg:1966:SKA

[Wei66a] Alvin Weinberg. Samuel K. Allison, 1901–1965. Bulletin of the Atomic Scientists, 22(1):2, January 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1966:CTR

[Wei66b] Alvin M. Weinberg. Can technology replace social engineering? Bulletin of the Atomic Scientists, 22(10):4–8, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Spi67].

Weinberg:1966:SCH

[Wei66c] Alvin M. Weinberg. Science, choice, and human values. Bulletin of the Atomic Scientists, 22(4):8–13, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weir:1966:UP

[Wei66d] John M. Weir. The unconquered plague. Bulletin of the Atomic Scientists, 22(8):46–48, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weidenbaum:1968:MPD

[Wei68a] Murray L. Weidenbaum. A matter for the public to decide? Bulletin of the Atomic Scientists, 24(6):7, June 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1968:LUP

[Wei68b] Alvin M. Weinberg. Let us prepare for peace. Bulletin of the Atomic Scientists, 24(7):17–20, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1968:STB

[Wei68c] V. F. Weisskopf. Spearhead toward a better world. Bulletin of the Atomic Scientists, 24(3):13, March 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from the author’s article Tribute to Marie Curie, Cern Courier, November 1967.

Weinberg:1969:TIC

[Wei69a] Alvin M. Weinberg. The Third International Conference on Science and Society. Bulletin of the Atomic Scientists, 25(9):23–26, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:1969:CDS

[Wei69b] Edward S. Weiss. Comments: Disputes Sternglass statistics. Bulletin of the Atomic Scientists, 25(8):28, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e, Sag69, Wor69, Ste69c].

Weidenbaum:1970:HBC

[Wei70a] Murray L. Weidenbaum. How to buy a cleaner environment. Bulletin of the Atomic Scientists, 26(9):19–21, November 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1970:PAR

[Wei70b] Alvin M. Weinberg. Part 3: Application and research: Nuclear energy and the environment. Bulletin of the Atomic Scientists, 26 (6):69–74, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1971:FRD

[Wei71] Alvin M. Weinberg. A far-reaching decision. Bulletin of the Atomic Scientists, 27(6):29, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1975:LCE

[Wei75] Victor Weisskopf. Letters: A clerical error. Bulletin of the Atomic Scientists, 31(4):3, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reports incorrect inclusion of Weisskopf’s name in [BAA^{+}75].

Weinberg:1976:WMR

[Wei76a] Alvin Weinberg. We must restructure the nuclear enterprise. Bulletin of the Atomic Scientists, 32(9):29, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Goo77, Rob77].

Weinberg:1976:MDS

[Wei76b] Alvin M. Weinberg. The many dimensions of scientific responsibility. Bulletin of the Atomic Scientists, 32(9):21–25, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Har77b, Mul77, Wal77, Wei77b].

Weinberg:1977:NEA

[Wei77a] Alvin M. Weinberg. Is nuclear energy acceptable? Bulletin of the Atomic Scientists, 33(4):54–60, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [MW77].

Weinberg:1977:LLH

[Wei77b] Alvin M. Weinberg. Letter: The linear hypothesis. Bulletin of the Atomic Scientists, 33(5):3–4, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei76b, Mul77, BW77].

Weiler:1978:CWR

[Wei78a] Lawrence D. Weiler. Commentary: From Washington: Reflections on the Disarmament Session. Bulletin of the Atomic Scientists, 34(10):7–9, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiner:1978:RSR

[Wei78b] Charles Weiner. Retroactive saber rattling? A note on nuclear physics in Japan. Bulletin of the Atomic Scientists, 34(4):10–12, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinstein:1978:CUS

[Wei78c] Franklin B. Weinstein. Commentary: The United States and the security of Southeast Asia. Bulletin of the Atomic Scientists, 34 (10):26–32, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinstein:1978:MNS

[Wei78d] Franklin B. Weinstein. The meaning of national security in Southeast Asia. Bulletin of the Atomic Scientists, 34(9):20–28, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:1978:NNP

[Wei78e] Leonard Weiss. The Nuclear Non-Proliferation Act of 1978. Bulletin of the Atomic Scientists, 34(3):31, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:1978:NSC

[Wei78f] Leonard Weiss. Nuclear safeguards: a Congressional perspective. Bulletin of the Atomic Scientists, 34(3):27–33, March 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Lov78a, Hur78b].

Weizenbaum:1978:OMC

[Wei78g] Joseph Weizenbaum. Once more — a computer revolution. Bulletin of the Atomic Scientists, 34(7):12–19, September 1978. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1979:PH

[Wei79] Victor F. Weisskopf. A peril and a hope. Bulletin of the Atomic Scientists, 35(1):10–13, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Win79].

Weinberg:1980:NEN

[Wei80a] Alvin M. Weinberg. Is nuclear energy necessary? Bulletin of the Atomic Scientists, 36(3):31–35, March 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BW80, Tho80a].

Weisskopf:1980:DES

[Wei80b] Victor F. Weisskopf. The double-edged sword called technology. Bulletin of the Atomic Scientists, 36(4):17–21, April 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1980:POP

[Wei80c] Victor F. Weisskopf. Perspective: The overwhelming priority. Bulletin of the Atomic Scientists, 36(2):1, February 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1982:AET

[Wei82a] Alvin M. Weinberg. Avoiding the entropy trap. Bulletin of the Atomic Scientists, 38(8):32–35, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MW83].

Weisskopf:1982:PNW

[Wei82b] Victor Weisskopf. Perspective: Nuclear war: four pressure points. Bulletin of the Atomic Scientists, 38(1):2, January 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weil:1983:SUC

[Wei83a] Irwin Weil. The Soviet Union: The cultural scene. Bulletin of the Atomic Scientists, 39(3):28–33, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiler:1983:NFU

[Wei83b] Lawrence D. Weiler. No first use: a history. Bulletin of the Atomic Scientists, 39(2):28–34, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1983:AAW

[Wei83c] Victor F. Weisskopf. Los Alamos anniversary: “We meant so well”. Bulletin of the Atomic Scientists, 39(7):24–26, August/ September 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weissbourd:1984:NWO

[Wei84] Bernard Weissbourd. Are nuclear weapons obsolete? Bulletin of the Atomic Scientists, 40(7):8–12, August/September 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ric84].

Weinberg:1985:SH

[Wei85a] Alvin M. Weinberg. The sanctification of Hiroshima. Bulletin of the Atomic Scientists, 41(11):34, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [NK86a].

Weinberg:1985:BRB

[Wei85b] Steve Weinberg. Book review: The Government/Press Connection: Press Officers and their Offices, by Stephen Hess. Bulletin of the Atomic Scientists, 41(3):60, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1985:LBA

[Wei85c] Victor F. Weisskopf. Looking back on Los Alamos. Bulletin of the Atomic Scientists, 41(7):20–22, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pru85].

Weinberg:1986:CESb

[Wei86a] Alvin M. Weinberg. Chernobyl: the emerging story: A nuclear power advocate reflects on Chernobyl. Bulletin of the Atomic Scientists, 42(7):57–60, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste86b].

Weinberg:1986:CESa

[Wei86b] Steve Weinberg. Chernobyl: the emerging story: Armand Hammer’s unique diplomacy. Bulletin of the Atomic Scientists, 42(7): 50–52, August/September 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weingarten:1986:LSR

[Wei86c] Lee Weingarten. Letter: Soviet ruling class. Bulletin of the Atomic Scientists, 42(1):56, January 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lon85c].

Weiss:1986:C

[Wei86d] Leonard Weiss. A commentary. Bulletin of the Atomic Scientists, 42(6):27–28, June/July 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1986:MOP

[Wei86e] Victor F. Weisskopf. In memoriam — Olof Palme. Bulletin of the Atomic Scientists, 42(5):49, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1987:TNP

[Wei87] Victor F. Weisskopf. The task for a new peace movement. Bulletin of the Atomic Scientists, 43(1):26–32, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weickhardt:1988:CP

[Wei88] George G. Weickhardt. Capitalists for perestroika. Bulletin of the Atomic Scientists, 44(5):30–34, June 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1989:BRBa

[Wei89a] Steve Weinberg. Book review: On Bended Knee: The Press and the Reagan Presidency, by Mark Hertsgaard. Bulletin of the Atomic Scientists, 45(1):62–63, January/February 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Ste89f].

Weinberg:1989:BRBb

[Wei89b] Steve Weinberg. Book review: The Master Spy, by Philip Knightley. Bulletin of the Atomic Scientists, 45(9):49, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1991:SJS

[Wei91a] Steve Weinberg. Soviet journalists: Starting to dig. Bulletin of the Atomic Scientists, 47(6):22–25, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:1991:PTN

[Wei91b] Leonard Weiss. Perspective: Tighten up on nuclear cheaters. Bulletin of the Atomic Scientists, 47(4):11–12, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1993:LGH

[Wei93] Victor Weisskopf. Letter: Giving Heisenberg his due. Bulletin of the Atomic Scientists, 49(9):53, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Swe93].

Weiner:1994:LFD

[Wei94a] Nella Fermi Weiner. Letter: Fermi: Didn’t even tell his wife. Bulletin of the Atomic Scientists, 50(4):3, July/August 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisgall:1994:ABW

[Wei94b] Jonathan M. Weisgall. The Able-Baker-Where’s-Charlie follies. Bulletin of the Atomic Scientists, 50(3):26–34, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisgall:1994:GOT

[Wei94c] Jonathan M. Weisgall. Guest opinion: Time to end the 40year lie. Bulletin of the Atomic Scientists, 50(3):3, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Nor94b].

Weisman:1994:ERW

[Wei94d] Jonathan Weisman. Early retirement for weaponeers? Bulletin of the Atomic Scientists, 50(4):16–22, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://www.highbeam.com/doc/1G1-15587904.html.

Weisman:1994:NR

[Wei94e] Jonathan Weisman. Nuckolls rapped. Bulletin of the Atomic Scientists, 50(4):18–19, July/August 1994. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

www.highbeam.com/doc/1G1-15587906.html.

Weisman:1997:WMS

[Wei97] Jonathan Weisman. Who’s minding the store? Bulletin of the Atomic Scientists, 53(4):32–37, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:2003:AP

[Wei03] Leonard Weiss. Atoms for peace. Bulletin of the Atomic Scientists, 59(6):34–44, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/59/6/34.

Weiss:2004:PDV

[Wei04] Leonard Weiss. Pakistan: It’s Déjà vu all over again. Bulletin of the Atomic Scientists, 60(3):52–59, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/52.

Weiss:2006:PP

[Wei06] Leonard Weiss. Power points. Bulletin of the Atomic Scientists, 62(3):21, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/62/3/21.

Weiss:2013:LAH

[Wei13] Leonard Weiss. The Lavon Affair: How a false-flag operation led to war and the Israeli bomb. Bulletin of the Atomic Scientists, 69(4):58–68, July 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/69/4/58.

Weiss:2015:FNT

[Wei15] Leonard Weiss. On fear and nuclear terrorism. Bulletin of the Atomic Scientists, 71(2):75–87, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiss:2017:SNW

[Wei17] Leonard Weiss. Safeguards and the NPT: Where our current problems began. Bulletin of the Atomic Scientists, 73(5):328– 336, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2017.1362906.

Weiss:2018:MNP

[Wei18] Leonard Weiss. The making of a non-proliferation law: A memoir. Bulletin of the Atomic Scientists, 74(4):276–282, 2018. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2018.1486617.

Wells:1951:WF

[Wel51] H. G. Wells. A Wellsian forecast. Bulletin of the Atomic Scientists, 7(10):317, October 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wells:1960:OTD

[Wel60] Warner Wells. Our technological dilemma, or an appraisal of man as a species bent on self-destruction. Bulletin of the Atomic Scientists, 16(9):362–365, November 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wells:1977:LUW

[Wel77] Merle W. Wells. Letter: University weapons work. Bulletin of the Atomic Scientists, 33(8):3–4, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day77k].

Welch:1979:BRB

[Wel79] Bruce L. Welch. Book review: The Nugget File, edited by Robert D. Pollard. Bulletin of the Atomic Scientists, 35(8):36–37, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Welch:1980:DNP

[Wel80] Bruce L. Welch. Deception on nuclear power risks: a call for action. Bulletin of the Atomic Scientists, 36(7):50–54, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wellerstein:2008:IAP

[Wel08] Alex Wellerstein. Inside the atomic patent office. Bulletin of the Atomic Scientists, 64(2):26–31, May/June 2008. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/64/2/26.

Wellerstein:2016:PPN

[Wel16] Alex Wellerstein. The psychological power of nuclear weapons. Bulletin of the Atomic Scientists, 72(5):298–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wentzel:1953:ILN

[Wen53] A. L. Wentzel. International laboratory for nuclear research. Bulletin of the Atomic Scientists, 9(6):223–224, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wermuth:1956:RCR

[Wer56] Anthony L. Wermuth. Reader comment: Ready — but not eager. Bulletin of the Atomic Scientists, 12(7):278–279, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wershow:1973:CIC

[Wer73] Harold Wershow. Communications: International censorship. Bulletin of the Atomic Scientists, 29(2):2, February 1973. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Werleigh:1993:SHH

[Wer93] Claudette Antoine Werleigh. Sanctions: Haiti and the halfhearted. Bulletin of the Atomic Scientists, 49(9):20–23, November 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westergaard:1955:MRH

[Wes55] Mogens Westergaard. Man’s responsibility to his genetic heritage. Bulletin of the Atomic Scientists, 11(9):318–328, 352, November 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wesley:1958:LEC

[Wes58] David Wesley. Letter to the Editor: Case for a bomb test ban. Bulletin of the Atomic Scientists, 14(5):200, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ore58b].

Westing:1971:CRE

[Wes71] Arthur H. Westing. Communications: A request [ecological impact of bomb craters]. Bulletin of the Atomic Scientists, 27(4):4, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westing:1974:ACE

[Wes74a] Arthur H. Westing. Arms control and the environment: Proscription of ecocide. Bulletin of the Atomic Scientists, 30(1): 24–27, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westing:1974:LGG

[Wes74b] Arthur H. Westing. Letters: A good guide. Bulletin of the Atomic Scientists, 30(5):5, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fra74].

Westcott:1976:LSP

[Wes76] Joan C. Westcott. Letter: A selling problem. Bulletin of the Atomic Scientists, 32(4):4, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ben76b].

Westervelt:1979:BNT

[Wes79a] Donald R. Westervelt. On banning nuclear tests: Can cold logic replace cold feet? Bulletin of the Atomic Scientists, 35(2):60– 62, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westing:1979:LEH

[Wes79b] Arthur H. Westing. Letter: An exchange with Hanoi. Bulletin of the Atomic Scientists, 35(3):88, March 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westing:1981:CDM

[Wes81] Arthur H. Westing. Crop destruction as a means of war. Bulletin of the Atomic Scientists, 37(2):38–42, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [MW81].

Westing:1985:BCW

[Wes85] Arthur H. Westing. Ban chemical weapons in Europe. Bulletin of the Atomic Scientists, 41(5):17–19, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Westfahl:1991:LBE

[Wes91] Wesley B. Westfahl. Letter: Bulletin editors unfeeling. Bulletin of the Atomic Scientists, 47(4):47, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack91a].

Wessells:1997:CS

[Wes97] Mike Wessells. Child soldiers. Bulletin of the Atomic Scientists, 53(6):32–39, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wesolowsky:1998:RNP

[Wes98] Tony Wesolowsky. Report: Nuclear power: Sparring over Mochovce. Bulletin of the Atomic Scientists, 54(6):19–20, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wesolowsky:1999:NPT

[Wes99] Tony Wesolowsky. Nuclear power: Tough choices on Temelin. Bulletin of the Atomic Scientists, 55(3):16–17, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/55/3/16.

Wesolowsky:2000:TFF

[Wes00a] Tony Wesolowsky. Temelin: Fueling up fast. Bulletin of the Atomic Scientists, 56(5):15–16, September/October 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/5/15.

Westwick:2000:BON

[Wes00b] Peter J. Westwick. In the beginning: The origin of nuclear security. Bulletin of the Atomic Scientists, 56(6):43–49, November/ December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/56/
6/43.

Wesolowsky:2001:RB

[Wes01] Tony Wesolowsky. Risky business. Bulletin of the Atomic Scientists, 57(3):40–44, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/3/40.

Wesolowsky:2003:CRR

[Wes03] Tony Wesolowsky. Czech Republic: From Russia, with love. Bulletin of the Atomic Scientists, 59(1):15–17, January/February 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/1/15.

Wesolowsky:2004:GBP

[Wes04] Tony Wesolowsky. Germany: Buying power. Bulletin of the Atomic Scientists, 60(3):13–14, May/June 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/3/13.

Wigner:1968:CCW

[WF68] Eugene P. Wigner and Bernard T. Feld. Current comments: Wigner and Feld. Bulletin of the Atomic Scientists, 24(2):29, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel67b, BJH68].

Webber:1977:LMW

[WF77] Carroll Webber, Jr. and Bernard T. Feld. Letter: Missile whistles. Bulletin of the Atomic Scientists, 33(10):9, December 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel77d].

Weickhardt:1986:NPC

[WF86] George C. Weickhardt and James M. Finberg. New push for chemical weapons. Bulletin of the Atomic Scientists, 42(9):28– 33, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1990:HEN

[WF90] Robert H. Williams and Harold A. Feiveson. How to expand nuclear power without proliferation. Bulletin of the Atomic Scientists, 46(3):40–45, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Cam90].

Winner:2016:RCD

[WF16] Andrew C. Winner and Ryan W. French. Rip currents: The dangers of nuclear-armed submarine proliferation. Bulletin of the Atomic Scientists, 72(4):222–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1969:NAL

[WG69] Robert Rathbun Wilson and Edwin L. Goldwasser. National Accelerator Laboratory: Policy statement on human rights, March 15, 1968. Bulletin of the Atomic Scientists, 25(8):9, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walder:1982:CBC

[WG82] Joseph Walder and Harold Green. Commentary: That Progressive case again. Bulletin of the Atomic Scientists, 38(8):59–60, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gre82b].

Weinberg:1972:GEI

[WH72] Alvin M. Weinberg and R. Philip Hammond. Global effects of increased use of energy. Bulletin of the Atomic Scientists, 28 (3):5–8, 43–44, March 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bar73b].

Wales:1989:PAP

[WH89] Jane Wales and Morton H. Halperin. Perspective: Advice to the president: Don’t count on nuclear weapons. Bulletin of the Atomic Scientists, 45(2):7–8, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weeks:2000:ESF

[WH00] Jennifer Weeks and John P. Holdren. Energy’s secrets: Finding the balance. Bulletin of the Atomic Scientists, 56(2):20–21, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/2/20.

Walker:2011:LRF

[WH11] Paul F. Walker and Jonathan R. Hunt. The legacy of Reykjavik and the future of nuclear disarmament. Bulletin of the Atomic Scientists, 67(6):63–72, November/December 2011. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/67/6/63.

Weise:2016:FIR

[WH16] Rachel A. Weise and Gretchen E. Hund. Financial incentives for reducing proliferation risks. Bulletin of the Atomic Scientists, 72(5):332–??, 2016. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wheaton:1951:FAT

[Whe51] William L. C. Wheaton. Federal action toward a national dispersal policy. Bulletin of the Atomic Scientists, 7(9):271–275, September 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wheeler:1964:CBS

[Whe64] Harvey Wheeler. The challenge of “Bureaucratized science”. Bulletin of the Atomic Scientists, 20(1):14–17, January 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Whealdon:1984:CBB

[Whe84a] Everett Whealdon. Commentary: Big brain. Bulletin of the Atomic Scientists, 40(7):51, August/September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bau84].

Wheatley:1984:BRB

[Whe84b] Steven C. Wheatley. Book reviews: Philanthropy and Cultural Imperialism: The Foundations at Home and Abroad, edited by Robert Arnove; Private Poewr for the Public Good: A History of the Carnegie Foundation for the Advancement of Teaching, by Ellen Condliffe Lagemann. Bulletin of the Atomic Scientists, 40 (1):40–42, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wheeler:1990:LHP

[Whe90] Willard T. Wheeler. Letter: Helms promotes Cold War and tobacco. Bulletin of the Atomic Scientists, 46(5):42, June 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Isa90h].

White:1948:RP

[Whi48a] Stephen White. Report from Paris. Bulletin of the Atomic Scientists, 4(11):324–325, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1948:ROR

[Whi48b] Stephen White. Report on Oak Ridge hearings. Bulletin of the Atomic Scientists, 4(7):194–196, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1951:CE

[Whi51a] A. S. White. Chemical engineering. Bulletin of the Atomic Scientists, 7(7–8):232–233, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1951:LPU

[Whi51b] Leonard D. White. The loyalty program of the United States Government. Bulletin of the Atomic Scientists, 7(12):363–366, 382, December 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Whiting:1953:LEO

[Whi53] Franck T. Whiting. Letter to the Editor: Our good Inquisitor General in Congress. Bulletin of the Atomic Scientists, 9(9):353, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1964:VFC

[Whi64] Gilbert F. White. Vietnam: The fourth course: International cooperation in science. Bulletin of the Atomic Scientists, 20(10): 6–10, December 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

WHCACD:1966:ACDa

[Whi66a] White House Committee on Arms Control and Disarmament. Arms control and disarmament at the White House. Bulletin of the Atomic Scientists, 22(2):43–45, February 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano66i] for corrections to committee member names.

WHCACD:1966:ACDb

[Whi66b] White House Committee on Arms Control and Disarmament. Arms control and disarmament at the White House. Bulletin of the Atomic Scientists, 22(3):33–37, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1971:EPM

[Whi71] Irvin L. White. Energy policy-making: Limitations of a conceptual model. Bulletin of the Atomic Scientists, 27(8):20–26, October 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1972:CE

[Whi72] Park J. White. Communications: On ecotheology. Bulletin of the Atomic Scientists, 28(10):2, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com72, Bur72].

Whitehurst:1975:LMC

[Whi75] Don Whitehurst. Letters: Move the clock. Bulletin of the Atomic Scientists, 31(6):2, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

White:1976:LMF

[Whi76a] David A. White. Letter: Mining the flowers. Bulletin of the Atomic Scientists, 32(9):4, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [PW76a, Ant76].

Whitelaw:1976:LCD

[Whi76b] Aubrey H. Whitelaw. Letter: The common defense. Bulletin of the Atomic Scientists, 32(10):2, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel76i].

White:1977:LVG

[Whi77] David E. White. Letter: Violence glorified. Bulletin of the Atomic Scientists, 33(1):3, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Whitelaw:1978:LEF

[Whi78a] A. H. Whitelaw. Letter: Exercise in futility. Bulletin of the Atomic Scientists, 34(5):7, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Kap78b].

Whitelaw:1978:LSL

[Whi78b] A. H. Whitelaw. Letter: To save lives. Bulletin of the Atomic Scientists, 34(4):5, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Whitmore:1979:CWS

[Whi79] B. G. Whitmore. Commentary: World security system. Bulletin of the Atomic Scientists, 35(4):54–55, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lei78].

Whiting:1989:BRB

[Whi89] Allen S. Whiting. Book review: Enter the Dragon, by Russel Spurr. Bulletin of the Atomic Scientists, 45(6):44–45, July/ August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Whyte:1959:UVE

[Why59] Lancelot Law Whyte. The unity of visual experience. Bulletin of the Atomic Scientists, 15(2):72–75, February 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Whyte:1962:BBS

[Why62] Lancelot Law Whyte. Books: The Sciences and the Arts: A New Alliance, by H. G. Cassidy. Bulletin of the Atomic Scientists, 18 (7):34–35, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Whyte:1971:SOU

[Why71] Lancelot Law Whyte. Science and our understanding of ourselves. Bulletin of the Atomic Scientists, 27(3):32–33, March 1971. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1971:CMD

[WI71] Alvin M. Weinberg and David Rittenhouse Inglis. Communications: The Malthusian dilemma: Weinberg and Inglis. Bulletin of the Atomic Scientists, 27(6):3, 39–41, June 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ing71].

Widditsch:1960:NR

[Wid60] Ann Widditsch. News roundup. Bulletin of the Atomic Scientists, 16(10):418–420, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:MHHa

[Wid61a] Ann Widditsch. Man and his habitat: Part II. Bulletin of the Atomic Scientists, 17(9):370, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:MHHb

[Wid61b] Ann Widditsch. Man and his habitat: Part III. Bulletin of the Atomic Scientists, 17(10):415, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:NRa

[Wid61c] Ann Widditsch. News roundup. Bulletin of the Atomic Scientists, 17(1):38–41, January 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:NRb

[Wid61d] Ann Widditsch. News roundup. Bulletin of the Atomic Scientists, 17(2):78–80, February 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:NRc

[Wid61e] Ann Widditsch. News roundup. Bulletin of the Atomic Scientists, 17(3):117–120, March 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:NRd

[Wid61f] Ann Widditsch. News roundup. Bulletin of the Atomic Scientists, 17(4):163–166, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:NRe

[Wid61g] Ann Widditsch. News roundup. Bulletin of the Atomic Scientists, 17(5–6):251–254, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Widditsch:1961:PC

[Wid61h] Ann Widditsch. Project Chariot. Bulletin of the Atomic Scientists, 17(10):426–427, December 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiener:1947:LES

[Wie47a] Norbert Wiener. Letter to the Editor: A scientists rebels. Bulletin of the Atomic Scientists, 3(1):31, January 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiener:1947:RRA

[Wie47b] Norbert Wiener. Replies to Ridenour: The Armed Services are not fit almoners for research. Bulletin of the Atomic Scientists, 3 (8):228, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiener:1948:FDR

[Wie48] Norbert Wiener. Freedom demands responsibility — the scientist’s role in society: A rebellious scientist after two years. Bulletin of the Atomic Scientists, 4(11):338–339, November 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [DS47a].

Wiener:1956:MRM

[Wie56] Norbert Wiener. Moral reflections of a mathematician. Bulletin of the Atomic Scientists, 12(2):53–57, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiener:1957:WSM

[Wie57] Norbert Wiener. [What a scientist must know]. Bulletin of the Atomic Scientists, 13(5):180, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1962:FRS

[Wie62] Jerome B. Wiesner. The Federal role in science and technology. Bulletin of the Atomic Scientists, 18(9):42, 44–46, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1964:SPP

[Wie64] Jerome B. Wiesner. “Science in policy, policy in science. Bulletin of the Atomic Scientists, 20(2):36–40, February 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1967:CWD

[Wie67] Jerome B. Wiesner. The Cold War is dead, but the arms race rumbles on. Bulletin of the Atomic Scientists, 23(6):6–9, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1968:HG

[Wie68] Jerome B. Wiesner. Hope for GCD? Bulletin of the Atomic Scientists, 24(1):10–15, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1970:ACC

[Wie70] Jerome B. Wiesner. Arms control: Current prospects and problems. Bulletin of the Atomic Scientists, 26(5):6–8, 38–39, May 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiener:1978:LNL

[Wie78] Hesh Wiener. Letter: No laughing matter. Bulletin of the Atomic Scientists, 34(1):5, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Koe77].

Wiesner:1982:MS

[Wie82a] Jerome B. Wiesner. Is a moratorium safe? Bulletin of the Atomic Scientists, 38(7):6–7, August/September 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1982:WCT

[Wie82b] Jerome B. Wiesner. What is a Comprehensive Test Ban? Bulletin of the Atomic Scientists, 38(6):13, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesen:1983:CPO

[Wie83] David L. Wiesen. Commentary: Population obstacle. Bulletin of the Atomic Scientists, 39(1):50–51, January 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1984:BND

[Wie84] Jerome B. Wiesner. Breaking the nuclear deadlock: Unilateral confidence building. Bulletin of the Atomic Scientists, 40(1):45– 47, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1985:USM

[Wie85] Jerome B. Wiesner. The United States: A militarized society. Bulletin of the Atomic Scientists, 41(7):102–105, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesen:1987:LSB

[Wie87] David L. Wiesen. Letter: Stocking up for the big rerun. Bulletin of the Atomic Scientists, 43(6):59, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark87g].

Wier:2006:TJ

[Wie06] Anthony Wier. Traffick jamming. Bulletin of the Atomic Scientists, 62(5):72, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/72.

Wigner:1951:IDA

[Wig51] Eugene P. Wigner. Impact of the developments in atomic energy on the sciences. Bulletin of the Atomic Scientists, 7(3):66–69, 80, March 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1958:LED

[Wig58] Eugene P. Wigner. Letter to the Editor: Detection of nuclear weapons testing. Bulletin of the Atomic Scientists, 14(6):233–234, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1960:FCC

[Wig60] Eugene P. Wigner. Fallout: Criticism of a criticism. Bulletin of the Atomic Scientists, 16(3):107–108, March 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lap60a].

Wigner:1961:REW

[Wig61] Eugene P. Wigner. Recall the ends — while pondering means. Bulletin of the Atomic Scientists, 17(3):82–85, March 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kue61].

Wigner:1962:BBI

[Wig62a] Eugene Wigner. Books: The Inspiration of Science, by Sir George Thompson. Bulletin of the Atomic Scientists, 18(9):27– 28, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1962:BRB

[Wig62b] Eugene P. Wigner. Book review: American Scientists and Nuclear Weapons Policy, by Robert Gilpin. Bulletin of the Atomic Scientists, 18(8):29–30, October 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wigner:1966:CDW

[Wig66] Eugene P. Wigner. Civil defense: Wigner on Project Harbor. Bulletin of the Atomic Scientists, 22(2):21–22, February 1966. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wightman:1988:LPT

[Wig88] Colin Wightman. Letter: Program trading and SDI. Bulletin of the Atomic Scientists, 44(2):52, March 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1952:MDA

[Wil52] Gilbert Wilson. Moby Dick and the atom. Bulletin of the Atomic Scientists, 8(6):195–197, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiley:1953:CA

[Wil53a] Alexander Wiley. A call for action. Bulletin of the Atomic Scientists, 9(7):243, 284, September 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1953:BBL

[Wil53b] Jay Williams. Books: Education and Liberty, by James Bryan Conant. Bulletin of the Atomic Scientists, 9(4):147, 152, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1953:SD

[Wil53c] Ira Jewell Williams, Jr. Science and dogma. Bulletin of the Atomic Scientists, 9(6):219–220, July 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1953:AT

[Wil53d] Robert R. Wilson. Atomic taboos. Bulletin of the Atomic Scientists, 9(1):7, 12, February 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1959:SMF

[Wil59] Benjamin H. Williams. Scientific method in foreign policy. Bulletin of the Atomic Scientists, 15(10):419–421, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1962:LES

[Wil62] Theodore E. Wilson. Letter to the Editor: Suggests CD mail refusal. Bulletin of the Atomic Scientists, 18(1):37, January 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilcox:1966:OMP

[Wil66] G. Leslie Wilcox. Observations on medical practices. Bulletin of the Atomic Scientists, 22(6):51–56, June 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1967:CMB

[Wil67a] Lloyd Williams. Comments: On Max Born’s “Reflections”. Bulletin of the Atomic Scientists, 23(2):27–28, February 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bor65c].

Willrich:1967:RIC

[Wil67b] Mason Willrich. Reports: International control of nuclear power. Bulletin of the Atomic Scientists, 23(3):31–38, March 1967. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1967:CES

[Wil67c] Dick Wilson. China’s economic situation. Bulletin of the Atomic Scientists, 23(9):3–8, November 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilhoit:1968:CCE

[Wil68a] Francis M. Wilhoit. Current comments: The ethical dilemmas of mythic nationalism. Bulletin of the Atomic Scientists, 24(8): 44–45, October 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1968:BBF

[Wil68b] Jane Wilson. Books: Vanished, by Fletcher Knebel; The Cassiopeia Affair, by Chloe Zerwick and Harrison Brown. Bulletin of the Atomic Scientists, 24(7):28–29, September 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1968:CCU

[Wil68c] Jane Wilson. Current comments: Universities act on the Institute for Defense Analysis. Bulletin of the Atomic Scientists, 24(5):40, May 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1969:WCS

[Wil69a] Dick Wilson. Where China stands now: An introduction. Bulletin of the Atomic Scientists, 25(2):4–10, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Wilson:1969:BBO

[Wil69b] Jane Wilson. Books: Lawrence and Oppenheimer, by Nuell Pharr Davis. Bulletin of the Atomic Scientists, 25(1):31–32, January 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dav68, Dav86b].

Wilson:1969:BBN

[Wil69c] Jane Wilson. Books: The New Brahmins: Scientific Life in America, by Spencer Klaw. Bulletin of the Atomic Scientists, 25(3):35–36, March 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1970:P

[Wil70a] Jane Wilson. Prologue. Bulletin of the Atomic Scientists, 26 (6):2–3, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1970:CP

[Wil70b] Robert Rathbun Wilson. The conscience of a physicist. Bulletin of the Atomic Scientists, 26(6):30–34, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1971:BRS

[Wil71] Jane Wilson. Book reviews: Several lives and more: My World Line, by George Gamow, The Viking Press, New York, 1970, 178 pages; Enrico Fermi: Physicist, by Emilio Segrè, The University of Chicago Press, Chicago, 1970, 276 pages; My Several Lives, Memoirs of a Social Inventor, by James B. Conant, Harper & Row, New York, 1970, 701 pages. Bulletin of the Atomic Scientists, 27(2):47–48, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1972:RE

[Wil72a] Jane Wilson. A Russian experiment. Bulletin of the Atomic Scientists, 28(10):29–31, December 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1972:PPP

[Wil72b] Richard Wilson. Power policy — plan or panic? Bulletin of the Atomic Scientists, 28(5):29–30, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kor72].

Williams:1973:RIC

[Wil73a] Carroll M. Williams. The rise of the International Center of Insect Physiology and Ecology. Bulletin of the Atomic Scientists, 29(3): 11–14, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1973:TBC

[Wil73b] Carroll L. Wilson. Toward the biological control of pests: Activating a vision. Bulletin of the Atomic Scientists, 29(3):10, 14–16, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1973:NGB

[Wil73c] Richard Wilson. Natural gas is a beautiful thing? Bulletin of the Atomic Scientists, 29(7):35–40, September 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1974:AOT

[Wil74a] Jane Wilson. “All in our time”: Reminiscences of nuclear pioneers: An introduction. Bulletin of the Atomic Scientists, 30(4): 10–11, April 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1974:BBM

[Wil74b] Jane Wilson. Books: The Master of Light, by Dorothy Michelson Livingston. Bulletin of the Atomic Scientists, 30(5):53–54, May 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Dud75a].

Wilcox:1975:LNS

[Wil75a] Bertram F. Wilcox. Letters: No sacred cows. Bulletin of the Atomic Scientists, 31(7):2, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dun75].

Willrich:1975:TKP

[Wil75b] Mason Willrich. Terrorists keep out! the problem of safeguarding nuclear materials in a world of malfunctioning people. Bulletin of the Atomic Scientists, 31(5):12–16, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rea76].

Wilson:1975:AOTb

[Wil75c] Jane Wilson. ‘All in our time’: The end of youth and innocence. Bulletin of the Atomic Scientists, 31(6):12–14, June 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1975:AOTc

[Wil75d] Jane Wilson, editor. All in our Time. The reminiscences of Twelve Nuclear Pioneers. Bulletin of the Atomic Scientists, Chicago, IL, USA, 1975. 236 pp. LCCN ????

Wilson:1975:LCJ

[Wil75e] Richard Wilson. Letter: Crossing the jungle. Bulletin of the Atomic Scientists, 31(9):3, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1975:AOTa

[Wil75f] Robert R. Wilson. “All in our time”: a recruit for Los Alamos. Bulletin of the Atomic Scientists, 31(3):41–47, March 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1975:LGC

[Wil75g] Steven C. Wilson. Letters: Gnats and camels. Bulletin of the Atomic Scientists, 31(3):2, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [De 74].

Williamson:1976:LMS

[Wil76a] Virginia M. Williamson. Letter: Main street. Bulletin of the Atomic Scientists, 32(7):4, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Maw76].

Wilson:1976:BMB

[Wil76b] Jane Wilson. Books: Math beneath the ivory tower: Adventures of a Mathematician, by S. M. Ulam. Bulletin of the Atomic Scientists, 32(9):45–46, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilkes:1977:WCP

[Wil77a] Garrison Wilkes. The world’s crop plant germplasm — an endangered resource. Bulletin of the Atomic Scientists, 33(2):8–16, February 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Koc77, Sma77].

Wilson:1977:BPC

[Wil77b] Jane Wilson. Books: Physics comes of age: the lively autobiography of Philip M. Morse: In At the Beginnings: A Physicist’s Life. Bulletin of the Atomic Scientists, 33(9):46–47, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1977:HNP

[Wil77c] Richard Wilson. How to have nuclear power without weapons proliferation. Bulletin of the Atomic Scientists, 33(9):39–44, November 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Willens:1978:AJC

[Wil78a] Harold Willens. Advice to Jimmy Carter. Bulletin of the Atomic Scientists, 34(5):7–8, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1978:LMM

[Wil78b] Carroll L. Wilson. Letter: Military myths. Bulletin of the Atomic Scientists, 34(3):3–4, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1978:CCF

[Wil78c] Jane S. Wilson. Commentary: The Chinese at Fermilab: Accelerating friendship. Bulletin of the Atomic Scientists, 34(8): 62–63, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1979:RAF

[Wil79] Carroll L. Wilson. Reminiscences of the AEC’s first general manager: Nuclear energy: what went wrong. Bulletin of the Atomic Scientists, 35(6):13–17, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1980:BRBa

[Wil80a] Jane S. Wilson. Book review: The Uranium People, by Leona Marshall Libby. Bulletin of the Atomic Scientists, 36(1):53–54, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1980:BRBb

[Wil80b] Richard Wilson. Book review: Frozen Fire, by Lee Niedringhaus Davis. Bulletin of the Atomic Scientists, 36(7):39–40, 42, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1980:COP

[Wil80c] Richard Wilson. Commentary: On ‘oil power’. Bulletin of the Atomic Scientists, 36(6):65, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pre80].

Wilson:1980:SSN

[Wil80d] Richard Wilson. Soviet scientists on nuclear power. Bulletin of the Atomic Scientists, 36(2):44–46, February 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wildavsky:1981:BRB

[Wil81a] Aaron Wildavsky. Book review: Technological Risk: Its Perception and Handling in the European Community, edited by Mainholf Dierkes, Sam Edwards and Rob Coppock. Bulletin of the Atomic Scientists, 37(9):44–46, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1981:NLS

[Wil81b] Robert R. Wilson. The new literature of science. Bulletin of the Atomic Scientists, 37(4):1, 4, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1981:NI

[Wil81c] Ward Wilson. Nuclear ignorance. Bulletin of the Atomic Scientists, 37(9):24–25, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Willens:1982:CCF

[Wil82a] Harold Willens. Commentary: California freeze initiative. Bulletin of the Atomic Scientists, 38(5):64, May 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1982:RWC

[Wil82b] Robert Williams. Robert Williams comments. Bulletin of the Atomic Scientists, 38(4):49–50, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Red82].

Wilson:1982:CIP

[Wil82c] Richard Wilson. Commentary: India and Pakistan: the nuclear confrontation. Bulletin of the Atomic Scientists, 38(8):60–61, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Her83b].

Wilke:1983:TTB

[Wil83] John Wilke. Threshold Test Ban Treaty: Seismic verification. Bulletin of the Atomic Scientists, 39(3):4–5, March 1983. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Lei83a, Mar83d].

Wilson:1984:BRB

[Wil84] Andrew Wilson. Book reviews: The Alliance, by Richard J. Barnet; The End of World Order, by Richard Falk. Bulletin of the Atomic Scientists, 40(4):34–35, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Willrich:1985:NCM

[Wil85a] Mason Willrich. Nuclear crisis management. Bulletin of the Atomic Scientists, 41(3):50–52, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lon85a].

Wilson:1985:NBY

[Wil85b] Robert R. Wilson. Niels Bohr and the young scientists. Bulletin of the Atomic Scientists, 41(7):23–26, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1986:LDW

[Wil86a] Kevin Jon Williams. Letter: Dust wars. Bulletin of the Atomic Scientists, 42(8):56, October 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1986:LME

[Wil86b] Richard Wilson. Letter: Middle East nuclear plants. Bulletin of the Atomic Scientists, 42(5):47–48, May 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ram86b] and comments [Ram86a, Lov86].

Williams:1987:DBS

[Wil87] Roger Williams. The decline of the British science empire. Bulletin of the Atomic Scientists, 43(8):45–48, October 1987. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1989:LGE

[Wil89] Edwin L. Williams. Letter: Give ’em hell. Bulletin of the Atomic Scientists, 45(4):52, May 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1990:GER

[Wil90] Robert R. Wilson. Guest editorial: A rallying cry — renew the movement. Bulletin of the Atomic Scientists, 46(2):3, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Lac90].

Wilkinson:1992:BRL

[Wil92a] Gregg S. Wilkinson. Book review: Low-dose danger: Radiation-Induced Cancer from Low-Dose Exposure, by John Gofman. Bulletin of the Atomic Scientists, 48(1):43–44, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1992:ES

[Wil92b] Kinsey Wilson. Expensive shutdown. Bulletin of the Atomic Scientists, 48(5):46, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1992:LDR

[Wil92c] Kinsey Wilson. A last-ditch rescue effort. Bulletin of the Atomic Scientists, 48(5):43–46, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1992:LS

[Wil92d] Kinsey Wilson. Lights out for Shoreham. Bulletin of the Atomic Scientists, 48(5):40–46, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

William:1993:LMW

[Wil93] Craig William. Letter: Making waste. Bulletin of the Atomic Scientists, 49(6):54, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi93b].

Wilkinson:1995:LNN

[Wil95] Michael Wilkinson. Letter: Negating the NPT. Bulletin of the Atomic Scientists, 51(1):67, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Eps94].

Williams:1999:DG

[Wil99] Phil Williams. Drugs and guns. Bulletin of the Atomic Scientists, 55(1):46–48, January/February 1999. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/46.

Wilson:2015:WTN

[Wil15] Ward Wilson. Why are there no big nuke protests? Bulletin of the Atomic Scientists, 71(2):50–59, March 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winetrout:1953:WTI

[Win53] Kenneth Winetrout. What time is it? Bulletin of the Atomic Scientists, 9(2):58–60, March 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winn:1971:TND

[Win71] Ira Jay Winn. Toward a new definition of national security. Bulletin of the Atomic Scientists, 27(3):35–37, March 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winterberg:1979:CHP

[Win79] F. Winterberg. Commentary: A high price paid. Bulletin of the Atomic Scientists, 35(4):56, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wei79].

Winstanley:1981:BRB

[Win81a] Derek Winstanley. Book review: The Greenhouse Effect, by Harold W. Bernard, Jr. Bulletin of the Atomic Scientists, 37(8): 38–39, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winterberg:1981:CIC

[Win81b] Friedwardt Winterberg. Commentary: Inertial confinement. Bulletin of the Atomic Scientists, 37(10):57–58, December 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Don82].

Winkler:1984:YHC

[Win84a] Allan M. Winkler. A 40-year history of civil defense. Bulletin of the Atomic Scientists, 40(6):16–22, June/July 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winkler:1984:BRB

[Win84b] Allan M. Winkler. Book review: International Arms Control: Issues and Agreements, edited by Coit D. Blacker and Gloria Duffy.

Bulletin of the Atomic Scientists, 40(8):36, October 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winkler:1984:H

[Win84c] Allan M. Winkler. History. Bulletin of the Atomic Scientists, 40 (10):12s–13s [various between 20 and 21], December 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winsor:1987:LSW

[Win87] Roswitha A. Winsor. Letter: Star Wars ruse. Bulletin of the Atomic Scientists, 43(1):58–59, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bro86e].

Winkler:1990:BRT

[Win90] Allan M. Winkler. Book review: Thaw and the Third World From Kabul to Managua: Soviet–American Relations in the 1980s, by Fred Halliday. Bulletin of the Atomic Scientists, 46(3):48, 50, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winkler:1994:BRB

[Win94] Allan M. Winkler. Book review: One World or None: A History of the World Nuclear Disarmament Movement through 1953, by Lawrence S. Wittner. Bulletin of the Atomic Scientists, 50(1):60– 61, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winkler:1998:BRM

[Win98] Allan M. Winkler. Book review: Moved by the movement: Resisting the Bomb: A History of the World Nuclear Disarmament Movement, 1954–1970, by Lawrence S. Wittner. Bulletin of the Atomic Scientists, 54(4):59–60, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wionczek:1977:SQW

[Wio77] Miguel S. Wionczek. Some questions for the ‘world jamboree’. Bulletin of the Atomic Scientists, 33(10):29–32, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wionczek:1979:CMS

[Wio79a] Miguel S. Wionczek. Commentary from Mexico: Science and technology for development. Bulletin of the Atomic Scientists, 35 (4):45–48, April 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wionczek:1979:CVU

[Wio79b] Miguel S. Wionczek. Commentary from Vienna: UNCSTD was not a technical failure. Bulletin of the Atomic Scientists, 35 (10):50–53, December 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wionczek:1981:PGD

[Wio81] Miguel S. Wionczek. Perspective: Global detonators. Bulletin of the Atomic Scientists, 37(2):1, February 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wionczek:1983:PPA

[Wio83] Miguel Wionczek. Power plays in Asia. Bulletin of the Atomic Scientists, 39(3):13–17, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wionczek:1984:WUS

[Wio84] Miguel S. Wionczek. On: ‘Why the United States is in Central America’. Bulletin of the Atomic Scientists, 40(1):63–64, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ken83].

Wise:1955:LEL

[Wis55] Catherine Wise. Letter to the Editor: Long-range goals of world government. Bulletin of the Atomic Scientists, 11(5):201, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Withheld:1952:LESa

[Wit52a] [Name Withheld]. Letter to the Editor: [subscription cancellation]. Bulletin of the Atomic Scientists, 8(1):33, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Withheld:1952:LESb

[Wit52b] [Name Withheld]. Letter to the Editor: [subscription cancellation]. Bulletin of the Atomic Scientists, 8(1):33, January 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wit:1982:BTB

[Wit82] Joel S. Wit. Backing away from test ban treaties. Bulletin of the Atomic Scientists, 38(8):36–38, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Witherup:1995:LTW

[Wit95a] Bill Witherup. Letter: I, too, was ten years old. Bulletin of the Atomic Scientists, 51(4):3, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wittner:1995:BS

[Wit95b] Lawrence S. Wittner. Blacklisting Schweitzer. Bulletin of the Atomic Scientists, 51(3):55–61, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Withington:2001:AEA

[Wit01a] Thomas Withington. Afghanistan: The early anti-Taliban team. Bulletin of the Atomic Scientists, 57(6):13–15, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/6/13.

Withington:2001:ASI

[Wit01b] Thomas Withington. Arms sales: Israel and India partner up. Bulletin of the Atomic Scientists, 57(1):18–19, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/1/18.

Withington:2001:EAT

[Wit01c] Thomas Withington. The early anti-Taliban team. Bulletin of the Atomic Scientists, 57(6):13, November/December 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Withington:2001:RT

[Wit01d] Thomas Withington. The rise of the Taliban. Bulletin of the Atomic Scientists, 57(6):69–70, November 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/6/69.

Withington:2003:CDM

[Wit03a] Thomas Withington. Civil defense: March, citizens, March. Bulletin of the Atomic Scientists, 59(6):15–16, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/59/6/15.

Withington:2003:TSS

[Wit03b] Thomas Withington. Terrorism: Stung by stingers. Bulletin of the Atomic Scientists, 59(3):16–17, May/June 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/3/16.

Wittner:2004:PP

[Wit04] Lawrence S. Wittner. The power of protest. Bulletin of the Atomic Scientists, 60(4):20–26, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/4/20.

Wexler:1961:MS

[WJ61] H. Wexler and D. S. Johnson. Meteorological satellites. Bulletin of the Atomic Scientists, 17(5–6):185–190, May/June 1961. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wilson:1982:COR

[WK82] Richard Wilson and Lars Kristoferson. Commentary: Osirak raid discussion continues. Bulletin of the Atomic Scientists, 38(2):45, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ne’81].

Wang:2023:FTL

[WKK23] Maya Wang, Frederike Kaltheuner, and Amanda Klasing. The future of technology: Lessons from China. Bulletin of the Atomic Scientists, 79(3):170–173, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2023.
2199595.

Wallace:1957:TDS

[WKW57] Anthony F. C. Wallace, Lewis M. Killian, and Harry B. Williams. Three disaster studies: Reply to Philip Wylie. Bulletin of the Atomic Scientists, 13(4):145–146, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wyl56, Wyl57].

Walker:1984:PNT

[WL84] William Walker and Måns Lönnroth. Proliferation and nuclear trade: a look ahead. Bulletin of the Atomic Scientists, 40(4): 29–33, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Rot84].

Wilson:2021:IUF

[WL21] Henrietta Wilson and Filippa Lentzos. Introduction: UNSCOM and the future of WMD verification. Bulletin of the Atomic Scientists, 77(4):159–162, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2021.1941532.

Whitmore:1997:LIS

[WLP97] Don Whitmore, George Lewis, and Theodore Postol. Letter: It still won’t work. Bulletin of the Atomic Scientists, 53(5):58–60, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [LP97].

Wilkes:1978:SNP

[WM78] Owen Wilkes and Robert Mann. The story of Nukey Poo: the Navy shuts down and dismantles a nuclear reactor in Antarctica; and the dirt ends up in Georgia and California. Bulletin of the Atomic Scientists, 34(8):32–36, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Way:1982:CBW

[WM82] Katharine Way and Dexter Masters. Commentary: One World or None. Bulletin of the Atomic Scientists, 38(2):49, February 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mag81].

Woods:1984:CMM

[WM84] John R. Woods and Tamara K. Manning. Commentary: ‘Minutes to midnight’. Bulletin of the Atomic Scientists, 40(4):46–48, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodwell:1979:CCD

[WMRK79] George M. Woodwell, Gordon J. MacDonald, Roger Revelle, and C. David Kelling. Commentary: The carbon dioxide report. Bulletin of the Atomic Scientists, 35(8):56–57, October 1979. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wiesner:1993:EO

[WMT93] Jerome B. Wiesner, Philip Morrison, and Kosta Tsipis. Ending overkill. Bulletin of the Atomic Scientists, 49(2):12–23, March 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walt:1979:CTG

[WN79] Lewis W. Walt and Lauris Norstad. Commentary: Two generals on SALT. Bulletin of the Atomic Scientists, 35(8):62–63, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wohl:1954:TAR

[Woh54] R. Richard Wohl. Technical assistance — retrospect and review. Bulletin of the Atomic Scientists, 10(3):75–80, March 1954. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wohlstetter:1964:TPD

[Woh64] Albert Wohlstetter. Technology, prediction, and disorder. Bulletin of the Atomic Scientists, 20(8):11–15, October 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wohlstetter:1968:PNE

[Woh68] Albert Wohlstetter. Perspective on nuclear energy. Bulletin of the Atomic Scientists, 24(4):2–5, April 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wohlstetter:1978:PEA

[Woh78] Albert Wohlstetter. From proof of evidence of Albert Wohlstetter on behalf of Friends of the Earth Ltd. in the Windscale Inquiry, September 5–6, 1977. Bulletin of the Atomic Scientists, 34(6):49, June 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfe:1950:HWS

[Wol50] Hugh Wolfe. Hugh Wolfe’s statement. Bulletin of the Atomic Scientists, 6(4):98, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolke:1955:BBT

[Wol55] R. L. Wolke. Books: Isotopic Tracers, A Theoretical and Practical Manual for Biological Students and Research Workers, by G. E. Francis, W. Mulligan, and A. Wormall. Bulletin of the Atomic Scientists, 11(2):64, February 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfe:1958:GBP

[Wol58a] Hugh C. Wolfe. George Braxton Pegram, 1876–1958. Bulletin of the Atomic Scientists, 14(8):328, October 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfle:1958:EW

[Wol58b] Dael Wolfle. Educational waste. Bulletin of the Atomic Scientists, 14(9):364–367, November 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfe:1962:BBL

[Wol62] Robert Paul Wolfe. Books: The Limits of Defense, by Arthur Waskow. Bulletin of the Atomic Scientists, 18(6):36–37, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolff:1963:BBP

[Wol63] Robert P. Wolff. Books: Strategy, Politics, and Defense Budgets, by Warner Schilling, Paul Hammond, and Glenn Snyder. Bulletin of the Atomic Scientists, 19(2):31–33, February 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolf:1965:MHH

[Wol65] Leonard Wolf. Man and his habitat: Problems of water pollution: Cleaning up the Merrimack. Bulletin of the Atomic Scientists, 21 (4):16–22, April 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfe:1968:CCI

[Wol68] Richard R. Wolfe. Current comments: Individual participation in governmental decisions. Bulletin of the Atomic Scientists, 24(10): 32–, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolf:1971:CMI

[Wol71a] Charles Wolf, Jr. Communications: Military–industrial complexities. Bulletin of the Atomic Scientists, 27(9):2, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolf:1971:MIC

[Wol71b] Charles Wolf, Jr. Military–industrial complexities. Bulletin of the Atomic Scientists, 27(2):19–22, February 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolochow:1975:LLO

[Wol75] Aleck Wolochow. Letter: Lost opportunity. Bulletin of the Atomic Scientists, 31(10):2, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolf:1976:PSE

[Wol76] Martin Wolf. Photovoltaic solar energy conversion. Bulletin of the Atomic Scientists, 32(4):26–33, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Tri77].

Wolf:1978:CLN

[Wol78] Sidney M. Wolf. Commentary: Liquified natural gas. Bulletin of the Atomic Scientists, 34(10):20–25, December 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sch79a].

Wolfe:1980:CAN

[Wol80] Bertram Wolfe. Could America’s nuclear policies be counterproductive. Bulletin of the Atomic Scientists, 36(1):43–48, January 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfe:1982:CDM

[Wol82] James R. Wolfe. Commentary: Defense mechanisms. Bulletin of the Atomic Scientists, 38(4):53, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfenstein:1986:LMR

[Wol86] Lincoln Wolfenstein. Letter: Military research at Brookhaven. Bulletin of the Atomic Scientists, 42(5):46–47, May 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sam86].

Wollman:1988:LPT

[Wol88] Garrett A. Wollman. Letter: PCs to test glasnost. Bulletin of the Atomic Scientists, 44(6):53, July/August 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ML88].

Wolfenstein:1991:PEN

[Wol91a] Lincoln Wolfenstein. Perspective: End nuclear addiction. Bulletin of the Atomic Scientists, 47(4):13–15, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolford:1991:LWR

[Wol91b] Douglas L. Wolford. Letter: Will the real Robert M. White …. Bulletin of the Atomic Scientists, 47(6):45, July/August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ano90-50].

Wolfsthal:1993:RMI

[Wol93] Jon Brook Wolfsthal. Report: Missiles: The Israeli initiative. Bulletin of the Atomic Scientists, 49(7):8–9, September 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfsthal:2001:RSN

[Wol01a] Jon B. Wolfsthal. Russia: Surveying the nuclear cities. Bulletin of the Atomic Scientists, 57(4):15–17, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/4/15.

Wolfsthal:2001:SNC

[Wol01b] Jon B. Wolfsthal. Surveying the nuclear cities. Bulletin of the Atomic Scientists, 57(4):15, July/August 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wolfsthal:2017:PMV

[Wol17] Jon Wolfsthal. The political and military vulnerability of America’s land-based nuclear missiles. Bulletin of the Atomic Scientists, 73(3):150–153, 2017. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2017.1314996.

Woodward:1955:MCD

[Woo55] Sir Llewellyn Woodward. “By man came death”. Bulletin of the Atomic Scientists, 11(1):19–23, 34, January 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodward:1956:SRB

[Woo56] Llewellyn Woodward. Science and the relations between states. Bulletin of the Atomic Scientists, 12(4):119–124, April 1956. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woollard:1961:SRE

[Woo61] G. P. Woollard. Space research and Earth sciences. Bulletin of the Atomic Scientists, 17(5–6):175–180, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wooster:1962:LEP

[Woo62] Harold Wooster. Letter to the Editor: The publication explosion. Bulletin of the Atomic Scientists, 18(6):44, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [MBD^{+}62].

Woodwell:1965:BBN

[Woo65] George M. Woodwell. Books: Man and Nature in America, by Arthur A. Ekirch, Jr. Bulletin of the Atomic Scientists, 21(6): 54–55, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woods:1973:CRW

[Woo73a] F. W. Woods. Communications: Radioactive wastes on ice. Bulletin of the Atomic Scientists, 29(9):3, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, ZSA73c].

Wooster:1973:BBP

[Woo73b] Warren S. Wooster. Books: The Politics of the Ocean, by Edward Wenk, Jr. Bulletin of the Atomic Scientists, 29(4):52–53, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woods:1978:LRO

[Woo78] John R. Woods. Letter: A ray of optimism. Bulletin of the Atomic Scientists, 34(3):4, 6, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodward:1981:COP

[Woo81] Herbert N. Woodward. Commentary: Optimum population. Bulletin of the Atomic Scientists, 37(6):61–62, June/July 1981. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodward:1982:BRB

[Woo82] Beverly Woodward. Book reviews: The National Interest and the Human Interest: an Analysis of U.S. Foreign Policy, by Robert Johansen; The Politics of Human Rights, by Paula Newberg. Bulletin of the Atomic Scientists, 38(2):35–37, February 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woods:1985:LWE

[Woo85] John R. Woods. Letter: Write to the editor. Bulletin of the Atomic Scientists, 41(3):64, March 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woods:1986:LBB

[Woo86] John R. Woods. Letter: Budget balancing and nuclear war. Bulletin of the Atomic Scientists, 42(3):56, March 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wood:1991:NWO

[Woo91] Bernard Wood. New world order: A thin fabric. Bulletin of the Atomic Scientists, 47(5):31–32, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodruff:1992:BRW

[Woo92] Roy D. Woodruff. Book review: “We need to do something”: Cardinal Choices: Presidential Science Advising from the Atomic Bomb to SDI, by Gregg Herken. Bulletin of the Atomic Scientists, 48(6):44–46, July/August 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodward:1993:SYD

[Woo93] Susan L. Woodward. Sanctions: Yugoslavia: Divide and fail. Bulletin of the Atomic Scientists, 49(9):24–27, November 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1996:FS

[Woo96a] Colin Woodard. Fighting for the scraps. Bulletin of the Atomic Scientists, 52(3):56–59, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1996:NE

[Woo96b] Colin Woodard. Nuclear expansion. Bulletin of the Atomic Scientists, 52(3):58, May/June 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1996:RBM

[Woo96c] Colin Woodard. Report: The Balkans: Morning in Dubrovnik. Bulletin of the Atomic Scientists, 52(6):11–13, November/ December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1997:RBF

[Woo97a] Colin Woodard. Report: Bosnia: A fictional state of being. Bulletin of the Atomic Scientists, 53(1):9–11, January/February 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1997:RBB

[Woo97b] Colin Woodard. Report: Bulgaria: From bad to worse. Bulletin of the Atomic Scientists, 53(3):15–16, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1997:RGA

[Woo97c] Colin Woodard. Report: Guatemala: After the killing. Bulletin of the Atomic Scientists, 53(4):9–11, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1997:RNP

[Woo97d] Colin Woodard. Report: Nuclear power: Yankee yanked. Bulletin of the Atomic Scientists, 53(6):10–12, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1997:RPL

[Woo97e] Colin Woodard. Report: Panama: Life after the Zone. Bulletin of the Atomic Scientists, 53(5):12–14, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1997:RRC

[Woo97f] Colin Woodard. Report: Romania: Of cabbages not kings. Bulletin of the Atomic Scientists, 53(2):11–13, March/April 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1998:RBY

[Woo98a] Colin Woodard. Report: Bosnia: Your tax dollars at work. Bulletin of the Atomic Scientists, 54(2):9–11, March/April 1998. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1998:RKD

[Woo98b] Colin Woodard. Report: Korea: DMZ holiday. Bulletin of the Atomic Scientists, 54(3):12–14, May/June 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1998:RMI

[Woo98c] Colin Woodard. Report: Marshall Islands: You can’t go home again. Bulletin of the Atomic Scientists, 54(5):10–12, September/ October 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1998:RML

[Woo98d] Colin Woodard. Report: Micronesia: Lifelines grow short. Bulletin of the Atomic Scientists, 54(6):16–18, November/December 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1998:RRO

[Woo98e] Colin Woodard. Report: Religion: An Orthodox odyssey. Bulletin of the Atomic Scientists, 54(1):9–11, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1998:RRG

[Woo98f] Colin Woodard. Report: Romania: Good neighbors. Bulletin of the Atomic Scientists, 54(4):13–15, July/August 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:1999:AED

[Woo99a] Colin Woodard. Antarctica: Endless détente. Bulletin of the Atomic Scientists, 55(1):10–12, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/1/10.

Woodard:1999:NTI

[Woo99b] Colin Woodard. NATO throws itself a party. Bulletin of the Atomic Scientists, 55(4):36–41, July/August 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/4/36.

Woodard:2000:BB

[Woo00a] Colin Woodard. Border brawl. Bulletin of the Atomic Scientists, 56(6):13, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:2000:BS

[Woo00b] Colin Woodard. Bosnia: Stillborn. Bulletin of the Atomic Scientists, 56(4):17–19, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/56/4/17.

Woodard:2000:CAB

[Woo00c] Colin Woodard. Central America: Border brawl. Bulletin of the Atomic Scientists, 56(6):13–15, November/December 2000. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/56/6/13.

Woodard:2000:MIP

[Woo00d] Colin Woodard. Marshall Islands: Payback time. Bulletin of the Atomic Scientists, 56(2):11–13, March 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/56/2/11.

Woodard:2000:MIT

[Woo00e] Colin Woodard. Marshall Islands: Time to re-up at Kwajalein. Bulletin of the Atomic Scientists, 56(1):19, January/February 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/1/19.

Woodard:2000:SBC

[Woo00f] Colin Woodard. Off-shore banking: Clean beaches, dirty money. Bulletin of the Atomic Scientists, 56(3):18–20, May/June 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/56/3/18.

Woodard:2000:S

[Woo00g] Colin Woodard. Stillborn. Bulletin of the Atomic Scientists, 56 (4):17, July/August 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Woodard:2001:DRM

[Woo01a] Colin Woodard. Decommissioning: Remember the Maine Yankee. Bulletin of the Atomic Scientists, 57(1):11–13, January/February 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/57/1/11.

Woodard:2001:LG

[Woo01b] Colin Woodard. A lackluster guide. Bulletin of the Atomic Scientists, 57(5):64–66, September/October 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/57/5/64.

Woodard:2001:SSO

[Woo01c] Colin Woodard. Space: Stuck in orbit. Bulletin of the Atomic Scientists, 57(2):11–13, March 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/57/2/11.

Woodard:2001:WCC

[Woo01d] Colin Woodard. War crimes: Called to account. Bulletin of the Atomic Scientists, 57(3):14–16, May/June 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/57/3/14.

Woodard:2002:ATT

[Woo02] Colin Woodard. Albania: a tale of two cities. Bulletin of the Atomic Scientists, 58(6):12–14, November 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/6/12.

Woodard:2003:KB

[Woo03] Colin Woodard. Kaliningrad: Boxed in. Bulletin of the Atomic Scientists, 59(5):16–18, September/October 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/59/5/16.

Woodard:2004:BBG

[Woo04] Colin Woodard. Bosnia: Back in the game. Bulletin of the Atomic Scientists, 60(4):12–14, July/August 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/60/4/12.

Woodard:2005:ASE

[Woo05] Colin Woodard. Albania: Southeast of Eden. Bulletin of the Atomic Scientists, 61(5):14–16, September/October 2005. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/61/5/14.

Wood:2012:PPS

[Woo12] Leet W. Wood. Projecting power: The security implications of space-based solar power. Bulletin of the Atomic Scientists, 68(1): 70–78, January/February 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/68/1/70.

Wornow:1958:LES

[Wor58] David Wornow. Letter to the Editor: Starving Russians? Bulletin of the Atomic Scientists, 14(5):201, May 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wormuth:1963:PB

[Wor63] Francis D. Wormuth. The politics of bedlam. Bulletin of the Atomic Scientists, 19(10):28–30, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Worth:1969:CCV

[Wor69] Dorothy J. Worth. Comments: “Confound that variable”. Bulletin of the Atomic Scientists, 25(8):28–29, October 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e, Sag69, Wei69b, Ste69c].

Worden:2006:HA

[Wor06] Simon Worden. High anxiety. Bulletin of the Atomic Scientists, 62(2):21–23, March/April 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/62/2/21.

Wouk:1971:ECB

[Wou71] Victor Wouk. Electric cars: The battery problem. Bulletin of the Atomic Scientists, 27(4):19–22, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wouk:1973:BBD

[Wou73] Victor Wouk. Books: The Death of the Automobile, by John Jerome. Bulletin of the Atomic Scientists, 29(9):54–56, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wouk:1976:LP

[Wou76] Herman Wouk. Letter: A protest. Bulletin of the Atomic Scientists, 32(9):7, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Wouk:1978:LNF

[Wou78] Victor Wouk. Letter: Nuclear fears exploded. Bulletin of the Atomic Scientists, 34(8):64, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sak78].

Wouk:1981:BRB

[Wou81] Victor Wouk. Book reviews: Running on Empty: The Future of the Automobile in an Oil Short World, by Lester R. Brown, Christopher Flavin and Colin Norman; Electric Automobiles: Energy Environmental and Economic Prospects for the Future, by William Hamilton. Bulletin of the Atomic Scientists, 37(7):48, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Winn:1980:CIE

[WP80] Ira J. Winn and Anthony Peranio. Commentary: Israel’s energy dilemma. Bulletin of the Atomic Scientists, 36(4):57–60, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weber:2019:SUM

[WP19] Andrew Weber and Christine Parthemore. Smarter US modernization, without new nuclear weapons. Bulletin of the Atomic Scientists, 75(1):25–29, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1555993.

Wolfe:1984:PC

[WPB84] Bertram Wolfe, George Palmer, and Dan Bolef. Plutonium connection. Bulletin of the Atomic Scientists, 40(7):50, August/ September 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [PB84].

Wright:1946:DCA

[Wri46] Quincy Wright. Draft for a convention on atomic energy. Bulletin of the Atomic Scientists, 1(8):11–13, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1947:BBN

[Wri47a] Quincy Wright. Books: Appeal to the Nations, by Norman Thomas. Bulletin of the Atomic Scientists, 3(7):195–196, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1947:UCP

[Wri47b] Quincy Wright. The U.N. Charter and the prevention of war. Bulletin of the Atomic Scientists, 3(2):57–58, 61, February 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1948:AIW

[Wri48] Quincy Wright. On the application of intelligence to world affairs. Bulletin of the Atomic Scientists, 4(8):249–252, August 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1949:CAG

[Wri49] Sewall Wright. Commentary by American geneticists: Dogma or opportunism? Bulletin of the Atomic Scientists, 5(5):141–142, May 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1961:UNS

[Wri61] Christopher Wright. United Nations and space. Bulletin of the Atomic Scientists, 17(5–6):236–240, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1963:BBD

[Wri63a] Quincy Wright. Books: NATO and the Defense of the West, by Prince Hubertus Zu Löwenstein and Volkmar Von Zühlsdorff, translated by Edward Fitzgerald; The Politics of Western Defense, by F. W. Mulley; British Foreign Policy since the Second World War, by C. M. Woodhouse; The Man on Horseback, the Role of the Military in Politics, by S. E. Finer. Bulletin of the Atomic Scientists, 19(7):33–34, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1963:UPP

[Wri63b] Robert L. Wright. U.S. patent policy and government research. Bulletin of the Atomic Scientists, 19(10):9–13, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1964:WP

[Wri64] Robert L. Wright. Wright on patents. Bulletin of the Atomic Scientists, 20(8):34, October 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1965:MIC

[Wri65] R. H. Wright. Metarchons: Insect control through recognition signals. Bulletin of the Atomic Scientists, 21(3):28–30, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1967:BBI

[Wri67] Quicy Wright. Books: Modern International Negotiation, Principles and Practice, by Arthur Lall. Bulletin of the Atomic Scientists, 23(8):33–34, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1977:RDT

[Wri77] Susan Wright. Recombinant DNA technology: who shall regulate? Bulletin of the Atomic Scientists, 33(8):4–5, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See rebuttal [BDE^{+}77].

Wright:1984:CT

[Wri84] R. H. Wright. Commentary: Thirsty? Bulletin of the Atomic Scientists, 40(3):62, March 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1985:MNB

[Wri85] Susan Wright. The military and the new biology. Bulletin of the Atomic Scientists, 41(5):10–16, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1987:NDB

[Wri87] Susan Wright. New designs for biological weapons. Bulletin of the Atomic Scientists, 43(1):43–46, January/February 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1989:CBW

[Wri89] Susan Wright. Chemical/biological weapons: The buildup that was. Bulletin of the Atomic Scientists, 45(1):52–56, January/ February 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1991:BTD

[Wri91] Susan Wright. Biowar treaty in danger. Bulletin of the Atomic Scientists, 47(7):36–40, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1997:RBC

[Wri97] Susan Wright. Report: Bioweapons: Cuba case tests treaty. Bulletin of the Atomic Scientists, 53(6):18–19, November/December 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1999:CNK

[Wri99a] David Wright. Cut North Korea some slack. Bulletin of the Atomic Scientists, 55(2):54–58, March 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/2/54.

Wright:1999:HU

[Wri99b] Susan Wright. The hijacking of UNSCOM. Bulletin of the Atomic Scientists, 55(3):23–25, May/June 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/3/23.

Wright:2002:BUV

[Wri02] Susan Wright. Bioweapons: U.S. vetoes verification. Bulletin of the Atomic Scientists, 58(2):24–26, March/April 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/58/2/24.

Wright:2004:TBT

[Wri04] Susan Wright. Taking biodefense too far. Bulletin of the Atomic Scientists, 60(6):58–66, November/December 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/60/6/58.

Wattenberg:1993:RB

[WRT93] Albert Wattenberg, Joseph Rotblat, and Kosta Tsipis. Remembering Bernie. Bulletin of the Atomic Scientists, 49(4):13–17, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weiner:1979:JRO

[WS79] Charles Weiner and Alice Kimball Smith. J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 35(1):57, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Request for additional Oppenheimer letters.

Wright:1983:RDB

[WS83] Susan Wright and Robert L. Sinsheimer. Recombinant DNA and biological warfare. Bulletin of the Atomic Scientists, 39(9):20–24, 26, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Kap83b].

Walker:1991:LDL

[WS91] Paul F. Walker and Eric Stambler. … and the dirty little weapons. Bulletin of the Atomic Scientists, 47(4):20–24, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Williams:1995:LCS

[WS95] Craig Williams and Amy E. Smithson. Letters: Chemical stew. Bulletin of the Atomic Scientists, 51(2):73–74, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi95b].

Wilson:2021:PUU

[WS21] Henrietta Wilson and Nikita Smidovich. Perspectives on UN-SCOM and UNMOVIC: An interview with Nikita Smidovich. Bulletin of the Atomic Scientists, 77(4):184–187, 2021. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/00963402.2021.1941564.

Waters:2015:CNW

[WSG^{+}15] Colin N. Waters, James P. M. Syvitski, Agnieszka Galuszka, Gary J. Hancock, Jan Zalasiewicz, Alejandro Cearreta, Jacques Grinevald, Catherine Jeandel, J. R. McNeill, Colin Summerhayes, and Anthony Barnosky. Can nuclear weapons fallout mark the beginning of the Anthropocene Epoch? Bulletin of the Atomic Scientists, 71(3):46–57, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wright:1984:RDB

[WSH^{+}84] Susan Wright, Robert L. Sinsheimer, Alexander Hiam, Michael Dortch, Michael E. Tate, and David Keppel. On ‘Recombinant DNA and biological warfare’. Bulletin of the Atomic Scientists, 40(2):59–63, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [WS83].

Wu:1960:RCM

[Wu60] Hsien Wu. Readers’ comments: A minimum program for peace. Bulletin of the Atomic Scientists, 16(5):185–186, May 1960. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wu:1999:PG

[Wu99] Changhua Wu. The price of growth. Bulletin of the Atomic Scientists, 55(5):58–66, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/5/58.

Wu:2022:WCW

[Wu22] (Clark) Aoqi Wu. What a Cold War crisis over Taiwan could tell us about China–Russia relations today. Bulletin of the Atomic Scientists, 78(5):261–267, 2022. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2022.
2109338.

Wu:2023:RTD

[Wu23] (Clark) Aoqi Wu. To reassure Taiwan and deter China, the United States should learn from history. Bulletin of the Atomic Scientists, 79(2):72–79, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2023.2178174.

Wunsch:1982:BRB

[Wun82] James S. Wunsch. Book reviews: Food for the Future, by Keith O. Campbell; New Roots for Agriculture, by Wes Jackson. Bulletin of the Atomic Scientists, 38(4):39–41, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurfel:1968:BBU

[Wur68] David Wurfel. Books: The United States in Vietnam, by George McT. Kahin and John W. Lewis. Bulletin of the Atomic Scientists, 24(2):21–23, February 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1991:BUC

[Wur91a] Jim Wurst. Bulletin: U.N. command of Gulf action unlikely. Bulletin of the Atomic Scientists, 47(1):4, January/February 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1991:RIU

[Wur91b] Jim Wurst. Report: Iraq: U.N. commission sifts through rubble. Bulletin of the Atomic Scientists, 47(7):10–12, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1992:RUN

[Wur92] Jim Wurst. Report: United Nations: A man, a plan, now what? Bulletin of the Atomic Scientists, 48(7):9–11, September 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1993:CM

[Wur93a] Jim Wurst. Clearing mines. Bulletin of the Atomic Scientists, 49(6):20, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1993:ELM

[Wur93b] Jim Wurst. Exposing the land mine business. Bulletin of the Atomic Scientists, 49(6):19, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1993:LMF

[Wur93c] Jim Wurst. The land mine family. Bulletin of the Atomic Scientists, 49(6):16, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1993:TMT

[Wur93d] Jim Wurst. Ten million tragedies, one step at a time. Bulletin of the Atomic Scientists, 49(6):14–21, July/August 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1994:MD

[Wur94a] Jim Wurst. Mozambique disarms. Bulletin of the Atomic Scientists, 50(5):36–39, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1994:MM

[Wur94b] Jim Wurst. Of mines and men. Bulletin of the Atomic Scientists, 50(5):39, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1994:RLM

[Wur94c] Jim Wurst. Report: Land mine: Still killing. Bulletin of the Atomic Scientists, 50(3):12–13, May/June 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1996:L

[Wur96a] Jim Wurst. Lasers. Bulletin of the Atomic Scientists, 52(2):11, March/April 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1996:RLMb

[Wur96b] Jim Wurst. Report: Land mines: Bobbled ban. Bulletin of the Atomic Scientists, 52(5):11–14, September/October 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1996:RLMa

[Wur96c] Jim Wurst. Report: Land mines: Inching toward a ban. Bulletin of the Atomic Scientists, 52(2):10–11, March/April 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1997:RMB

[Wur97] Jim Wurst. Report: Mozambique: Beating swords into … furniture. Bulletin of the Atomic Scientists, 53(3):17–18, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wurst:1999:UG

[Wur99] Jim Wurst. The U.N. gears up. Bulletin of the Atomic Scientists, 55(1):58–60, January/February 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/55/1/58.

Wolfe:1986:LLT

[WvH86] Bertram Wolfe and Frank von Hippel. Letter: Long-term health effects of Chernobyl. Bulletin of the Atomic Scientists, 42(9):55– 56, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vHC86].

Watson-Watt:1950:SPC

[WW50] Sir Robert Watson-Watt. Science, politics, and citizenship. Bulletin of the Atomic Scientists, 6(1):27–28, January 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Watson-Watt:1958:EF

[WW58] Sir Robert Watson-Watt. Emotional fall-out. Bulletin of the Atomic Scientists, 14(6):215–216, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Watson-Watt:1959:PP

[WW59] Robert Watson-Watt. Physicist and politics. Bulletin of the Atomic Scientists, 15(7):298–301, September 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weinberg:1960:LRV

[WW60] Alvin M. Weinberg and Eugene P. Wigner. Longer range view of nuclear energy. Bulletin of the Atomic Scientists, 16(10):400– 403, December 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mei61].

Weinberg:1961:LEW

[WW61] Alvin M. Weinberg and Eugene P. Wigner. Letter to the Editor: Why reactor development is economical. Bulletin of the Atomic Scientists, 17(4):120, April 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mei61].

Watson-Watt:1962:BBG

[WW62] Sir Robert Watson-Watt. Books: Sir George Cayley, The Inventor of the Aeroplane, by J. Laurence Pritchard. Bulletin of the Atomic Scientists, 18(7):31–32, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Weisskopf:1980:PAS

[WW80] Victor F. Weisskopf and Robert R. Wilson. Perspective: American–Soviet scientific exchanges: more important than ever. Bulletin of the Atomic Scientists, 36(6):1, June 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Walker:1986:MMS

[WW86] Paul F. Walker and John A. Wentworth. Midgetman: missile in search of a mission. Bulletin of the Atomic Scientists, 42(9):20– 26, November 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wittner:2007:AR

[WWHV07] Lawrence S. Wittner, Jessica Wilbanks, Kate Hudson, and Aaron Voldman. Activism revived. Bulletin of the Atomic Scientists, 63 (6):20–21, November/December 2007. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/63/6/20.

Wyart:1952:SFE

[Wya52] J. Wyart. Some French experiences. Bulletin of the Atomic Scientists, 8(7):240–241, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders..

Wylie:1951:IS

[Wyl51] Philip Wylie. Impure science. Bulletin of the Atomic Scientists, 7(7–8):196–200, August 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [HR51, deF51, Ma’52].

Wylie:1953:KMS

[Wyl53] Philip Wylie. “Knowledge for man’s sake”. Bulletin of the Atomic Scientists, 9(9):330–331, November 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Wylie:1954:PPB

[Wyl54] Philip Wylie. Panic, psychology, and the bomb. Bulletin of the Atomic Scientists, 10(2):37–40, 63, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See responses [Mar54c, Mer54b].

Wylie:1956:BBB

[Wyl56] Philip Wylie. Books: Human Behavior In Extreme Situations: A Survey of the Literature and Suggestions for Further Research, by Lewis M. Killian (with the assistance of Randolph Quick and Frank Stockwell); A Study of Response to the Houston, Texas, Fireworks Explosion; Tornado in Worcester. An Exploratory Study of Individual and Community Behavior in an Extreme Situation, by Anthony F. C. Wallace. Bulletin of the Atomic Scientists, 12(10):376–377, December 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [WKW57] and answer [Wyl57].

Wylie:1957:TDS

[Wyl57] Philip Wylie. Three disaster studies: Mr. Wylie’s answer. Bulletin of the Atomic Scientists, 13(4):146, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Wyl56, WKW57].

Wylie:1969:RWR

[Wyl69]  Ray Wylie. Revolution within a revolution?  Bulletin of

the Atomic Scientists, 25(2):29–32, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Weertman:1973:CRW

[WZSA73] J. Weertman, E. J. Zeller, D. F. Saunders, and E. E. Angino. Communications: Radioactive wastes on ice: Further discussion. Bulletin of the Atomic Scientists, 29(4):2–3, 53–54, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d].

Yaffa:2006:RSP

[Yaf06] Joshua Yaffa. Russia: Sparring partners. Bulletin of the Atomic Scientists, 62(4):18–19, July/August 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/4/18.

Yahuda:1969:CNO

[Yah69] Michael B. Yahuda. China’s nuclear option. Bulletin of the Atomic Scientists, 25(2):72–77, February 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special issue on China after the Cultural Revolution.

Yang:1982:CCN

[Yan82] Richard H. Yang. Commentary: China and the nuclear club. Bulletin of the Atomic Scientists, 38(4):54, April 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yarmolinsky:1980:BRB

[Yar80a] Adam Yarmolinsky. Book review: Can Organizations Change? Environmental Protection, Citizen Participation, and Corps of Engineers, by Daniel A. Mazmanian and Jeanne Neinaber. Bulletin of the Atomic Scientists, 36(7):46, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yarmolinsky:1980:CKT

[Yar80b] Adam Yarmolinsky. Commentary: Keeping tabs on R&D. Bulletin of the Atomic Scientists, 36(9):49–50, November 1980. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yarmolinsky:1980:CPT

[Yar80c] Adam Yarmolinsky. Commentary: Part two: Keeping tabs on R&D. Bulletin of the Atomic Scientists, 36(10):55–56, December 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yates:1969:CCS

[Yat69] Sidney R. Yates. Current comments: Showdown on the ABM. Bulletin of the Atomic Scientists, 25(3):29–32, March 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yazov:1988:PSP

[Yaz88] Dmitri Yazov. Perspective: The Soviet proposal for European security. Bulletin of the Atomic Scientists, 44(7):8–11, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yulish:1998:LMM

[YC98] Charles B. Yulish and Francesco Calogero. Letter: Megatons to megawatts. Bulletin of the Atomic Scientists, 54(1):3, 66– 67, January/February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Cal97].

Yemelyanov:1987:MSB

[Yem87] Vasily S. Yemelyanov. The making of the Soviet bomb. Bulletin of the Atomic Scientists, 43(10):39–41, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yeoman:1975:LNF

[Yeo75] Margaret Yeoman. Letters: Nuclear flack. Bulletin of the Atomic Scientists, 31(5):4, May 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi75b].

Yadav:2021:CMT

[YES^{+}21] Kamya Yadav, Ulas Erdogdu, Samikshya Siwakoti, Jacob N. Shapiro, and Alicia Wanless. Countries have more than 100 laws on the books to combat misinformation. How well do they work? Bulletin of the Atomic Scientists, 77(3):124–128, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2021.1912111.

Yevtushenko:1986:YYW

[Yev86] Yevgeny Yevtushenko. Yevgeny Yevtushenko on writing, censorship, and life in the Soviet Union. Bulletin of the Atomic Scientists, 42(2):11, February 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from the New York Times, 19 December 1985.

York:1977:MRD

[YG77a] Herbert F. York and G. Allen Greb. Military research and development: a postwar history: the changing role of scientists and engineers in shaping U.S. military technology programs. Bulletin of the Atomic Scientists, 33(1):12–22, 24–26, January 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1977:SR

[YG77b] Herbert F. York and G. Allen Greb. Strategic reconnaissance. Bulletin of the Atomic Scientists, 33(4):33–42, April 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yin:2015:HAN

[Yin15] Lu Yin. How to approach nuclear modernization? : a Chinese response. Bulletin of the Atomic Scientists, 71(3):8–11, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yoachim:1978:LRH

[Yoa78] John Yoachim. Letter: Rights of the Hopi. Bulletin of the Atomic Scientists, 34(1):7, January 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Hel77a].

Yondorf:1968:BBW

[Yon68] Walter F. Yondorf. Books: National Weapons Safety and the Common Defense, by Joel Larus. Bulletin of the Atomic Scientists, 24(9):38–39, November 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1969:MAM

[Yor69] Herbert F. York. Missiles and anti-missiles … six views: The arms race and the fallacy of the last move. Bulletin of the Atomic Scientists, 25(6):27–29, 43, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1970:CPV

[Yor70] Herbert F. York. Comments: A personal view of the arms race. Bulletin of the Atomic Scientists, 26(3):27–31, March 1970. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1973:CQA

[Yor73] Herbert F. York. Controlling the qualitative arms race. Bulletin of the Atomic Scientists, 29(3):4–8, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1974:DMM

[Yor74] Herbert F. York. Deterrence by means of mass destruction. Bulletin of the Atomic Scientists, 30(3):4–9, March 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Gar74, Kin74, Pow74].

York:1975:OLL

[Yor75a] Herbert F. York. The origins of the Lawrence Livermore Laboratory. Bulletin of the Atomic Scientists, 31(7):8–14, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1975:SAT

[Yor75b] Herbert F. York. Sounders of the alarm: Thirty years of attempts to bring the world to grips with the perpetual menace to human security. Bulletin of the Atomic Scientists, 31(10):43–45, December 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1976:NBT

[Yor76a] Herbert F. York. The nuclear ‘balance of terror’ in Europe. Bulletin of the Atomic Scientists, 32(5):8–14, 16, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1976:AOT

[Yor76b] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. W. H. Freeman, New York, NY, USA, 1976. ISBN 0-7167-0718-7. x + 175 pp. LCCN UG1282.A8 Y67.

York:1981:SNT

[Yor81] Herbert F. York. Sakharov and the nuclear test ban. Bulletin of the Atomic Scientists, 37(9):33–37, November 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Eva82].

York:1982:TSY

[Yor82] Herbert York. Thirty-seven years of nuclear weapons: Vertical proliferation. Bulletin of the Atomic Scientists, 38(10):47–50, December 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1983:BND

[Yor83] Herbert F. York. Breaking the nuclear deadlock: Negotiating from the bottom up. Bulletin of the Atomic Scientists, 39(8): 54–57, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1988:MWT

[Yor88] Herbert F. York. Making weapons, talking peace. Bulletin of the Atomic Scientists, 44(4):27–30, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [All88].

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67 1989. URL http://alsos.wlu.edu/

information.aspx?id=442;   http://www.loc.gov/catdir/
description/cam023/88062671.html;   http://www.loc.gov/
catdir/toc/cam027/88062671.html.

Yost:1980:NSR

[Yos80] Charles W. Yost. National security revisited. Bulletin of the Atomic Scientists, 36(8):20–24, October 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1947:RRS

[You47] Donald Young. Replies to Ridenour: Subsidies acceptable with proper safeguards. Bulletin of the Atomic Scientists, 3(8):229, August 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1956:BBC

[You56] Edwin Young. Books: A Chronicle of Jeopardy — 1945–55, by Rexford G. Tugwell. Bulletin of the Atomic Scientists, 12(3):93, March 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yout:1958:RMB

[You58] Rosalie C. Yout. Radiation and man: Basic data. Bulletin of the Atomic Scientists, 14(1):9–11, January 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1961:RGT

[You61] Wayland Young. Report from Geneva: Test ban negotiations. Bulletin of the Atomic Scientists, 17(5–6):250, May/June 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1962:MA

[You62] Wayland Young. Mosquitoes in Accra. Bulletin of the Atomic Scientists, 18(7):45–47, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1963:FDS

[You63a] Wayland Young. Forum: Disarmament steps. Bulletin of the Atomic Scientists, 19(7):32, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Pan63].

Young:1963:RLW

[You63b] Wayland Young. Report from London: The Western Alliance. Bulletin of the Atomic Scientists, 19(5):44, 46–48, May 1963. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1964:FMW

[You64a] Wayland Young. Forum: MLF — a West European view. Bulletin of the Atomic Scientists, 20(9):19–21, November 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1964:LEU

[You64b] Wayland Young. Letter to the Editor: U.S. China policy. Bulletin of the Atomic Scientists, 20(3):25, March 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1965:BBA

[You65a] Elizabeth Young. Books: The Arms Debate, by Robert A. Levine; The Dispersion of Nuclear Weapons: Strategy and Politics. Bulletin of the Atomic Scientists, 21(6):51–54, June 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1965:PAC

[You65b] Wayland Young. The prospects for arms control in Europe. Bulletin of the Atomic Scientists, 21(7):22–24, September 1965. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1966:BBB

[You66a] Elizabeth Young. Books: Must the Bomb Spread? by Leonard Beaton; Containing the Arms Race: Some Specific Proposals, by Jeremy J. Stone; Problems of National Strategy: A Book of Readings, by Henry A. Kissinger. Bulletin of the Atomic Scientists, 22(8):40–43, October 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1966:BBT

[You66b] Elizabeth Young. Books: Nato in Transition: The Future of the Atlantic Alliance, by Timothy M. Stanley. Bulletin of the Atomic Scientists, 22(4):37–38, April 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1966:BBP

[You66c] Elizabeth Young. Books: Strategic Power and Soviet Foreign Policy, by A. L. Horelick and Myron Rush; Khrushchev and the Arms Race: Soviet Interests in Arms Control and Disarmament, 1954– 1964, by Lincoln Bloomfield, Walter Clemens, Jr., and Franklyn Griffiths. Bulletin of the Atomic Scientists, 22(10):26–28, December 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1967:ANA

[You67a] Elizabeth Young. ABM: No alternatives to politics. Bulletin of the Atomic Scientists, 23(6):47–49, June 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1967:UUD

[You67b] Oran R. Young. The U.S. and the USSR debate the ABM active defense and international order. Bulletin of the Atomic Scientists, 23(5):34–42, May 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1968:BBC

[You68a] Elizabeth Young. Books: Arms Control for the Late Sixties, by James E. Dougherty and J. F. Lerman, Jr. Bulletin of the Atomic Scientists, 24(10):29–31, December 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1968:CCE

[You68b] Elizabeth Young. Current comments: Excerpts from BBC on ABM. Bulletin of the Atomic Scientists, 24(1):59, January 1968. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1968:PCA

[You68c] Oran R. Young. The political consequences of active defense. Bulletin of the Atomic Scientists, 24(2):16–20, February 1968. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1969:LED

[You69] Elizabeth Young. Letter to the Editor: Dyson may be right. Bulletin of the Atomic Scientists, 25(9):46, November 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys69a].

Young:1971:CPC

[You71] Irving G. Young. Communications: Population control. Bulletin of the Atomic Scientists, 27(9):2, November 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1973:CTV

[You73] Elizabeth Young. Communications: Tunnel vision. Bulletin of the Atomic Scientists, 29(3):2–3, March 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rat72].

Young:1975:RWS

[You75] Louise B. Young. Rhetoric won’t solve the safety problem. Bulletin of the Atomic Scientists, 31(7):52–54, September 1975. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [YY74, Sch75a, YY74].

Young:1976:BBB

[You76a] Elizabeth Young. Books: Britain and the bomb: Independence and Deterrence. Britain and Atomic Energy 1945–1952, by Margaret Gowing, assisted by Lorna Arnold. Bulletin of the Atomic Scientists, 32(8):46–47, October 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1976:LAI

[You76b] Elizabeth Young. Letter: Arms insurance. Bulletin of the Atomic Scientists, 32(7):6, September 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76d].

Young:1976:LCM

[You76c] Elizabeth Young. Letter: Cuban missiles. Bulletin of the Atomic Scientists, 32(5):3, May 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ber76b].

Young:1977:LRP

[You77a] Elizabeth Young. Letter: Russian PNEs. Bulletin of the Atomic Scientists, 33(1):4, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lon76a].

Young:1977:LWW

[You77b] Jean Young. Letter: Wasting wood. Bulletin of the Atomic Scientists, 33(4):4, 6, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Younghein:1977:LLF

[You77c] Ilene Younghein. Letter: A lonely feeling. Bulletin of the Atomic Scientists, 33(5):2–, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1978:CLS

[You78a] Elizabeth Young. Commentary: Law of the Sea — an arms control issue. Bulletin of the Atomic Scientists, 34(9):53–56, November 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1978:LNU

[You78b] Elizabeth Young. Letter: Nuclear umbrellas. Bulletin of the Atomic Scientists, 34(4):2, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste77b].

Younghein:1978:LRH

[You78c] Ilene Younghein. Letter: Radiation from heaven? Bulletin of the Atomic Scientists, 34(4):2–3, April 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1981:CEA

[You81] Elizabeth Young. Commentary: END appeal. Bulletin of the Atomic Scientists, 37(4):57, April 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [CY81].

Young:1982:BRB

[You82] Elizabeth Young. Book review: Between Peace and War: The Nature of International Crisis, by Richard Ned Lebow. Bulletin of the Atomic Scientists, 38(8):43–45, October 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1984:CMB

[You84a] Elizabeth Young. Commentary: On ‘The MX bargain’. Bulletin of the Atomic Scientists, 40(4):46, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Asp83, Tsi83b].

Young:1984:BNB

[You84b] Ruth Young. Book note: Trever’s First Strike, by John Trever. Bulletin of the Atomic Scientists, 40(6):40, June/July 1984. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1985:LST

[You85a] Elizabeth Young. Letter: Soviet treat violations. Bulletin of the Atomic Scientists, 41(6):54–55, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lon85b].

Young:1985:LPE

[You85b] Nigel Young. Letter: Peace education. Bulletin of the Atomic Scientists, 41(5):54, May 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rin84c].

Young:1987:LDC

[You87a] Elizabeth Young. Letter: Dispel conventional myths. Bulletin of the Atomic Scientists, 43(10):49–50, December 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ack87a].

Youngdale:1987:LSW

[You87b] James M. Youngdale. Letter: Star Wars culture. Bulletin of the Atomic Scientists, 43(2):61–62, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Smi87c].

Young:1991:NWO

[You91] Marilyn B. Young. New world order: Ruthless intervention. Bulletin of the Atomic Scientists, 47(5):32–33, June 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Young:1992:LEN

[You92] Elizabeth Young. Letter: Everyone’s nukes should head for trash. Bulletin of the Atomic Scientists, 48(3):47, April 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ADK91].

Yankoski:2021:MWA

[YSW21] Michael Yankoski, Walter Scheirer, and Tim Weninger. Meme warfare: AI countermeasures to disinformation should focus on popular, not perfect, fakes. Bulletin of the Atomic Scientists, 77 (3):119–123, 2021. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2021.1912093.

Yukawa:1961:MRS

[YTS61] Hideki Yukawa, Shimichiro Tomonaga, and Shoichi Sakata. Messages received at Stowe: From three Japanese scientists. Bulletin of the Atomic Scientists, 17(9):386, November 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yuan:2023:USS

[Yua23] Jingdong Yuan. The United States and stability in the Taiwan Strait. Bulletin of the Atomic Scientists, 79(2):80–86, 2023. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2023.2178176.

Yudken:1983:BRB

[Yud83] Joel S. Yudken. Book review: The Politics of Defense Contracting: The Iron Triangle, by Gordon Adams. Bulletin of the Atomic Scientists, 39(5):43–44, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yudelman:1985:ASS

[Yud85] Montague Yudelman. Africa: Sub-Saharan agricultural research. Bulletin of the Atomic Scientists, 41(8):35–38, September 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yukawa:1981:AE

[Yuk81] Hideki Yukawa. The absolute evil. Bulletin of the Atomic Scientists, 37(1):37, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yunusov:1994:AMM

[Yun94] Arif Yunusov. Azerbaijan: Malicious mapmaking. Bulletin of the Atomic Scientists, 50(1):25–27, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Yuter:1969:CA

[Yut69a] S. C. Yuter. Communications: An addendum. Bulletin of the Atomic Scientists, 25(10):43, December 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yut69b].

Yuter:1969:RWL

[Yut69b] S. C. Yuter. The role of world law in arms control. Bulletin of the Atomic Scientists, 25(8):23–25, October 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Yut69a].

Yankoski:2020:AEW

[YWS20] Michael Yankoski, Tim Weninger, and Walter Scheirer. An AI early warning system to monitor online disinformation, stop violence, and protect elections. Bulletin of the Atomic Scientists, 76(2):85–90, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2020.1728976.

Young:1974:PET

[YY74] Louise B. Young and H. Peyton Young. Pollution by electrical transmission: The environmental impact of high voltage lines. Bulletin of the Atomic Scientists, 30(10):34–38, December 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kor75, Sch75a, You75].

Zeman:2004:ACH

[ZA04] Scott C. Zeman and Michael A. Amundson, editors. Atomic culture: how we learned to stop worrying and love the bomb. Atomic history and culture. University Press of Colorado, Boulder, CO, USA, 2004. ISBN 0-87081-763-9 (hardcover), 0-87081-764-7 (paperback). x + 187 pp. LCCN E169.12 Z46 2004. URL http://

www.loc.gov/catdir/toc/ecip0413/2004001010.html.

Zacharias:1950:CVO

[Zac50] Jerrold R. Zacharias. A citizen’s view of our national security. Bulletin of the Atomic Scientists, 6(7):218–219, July 1950. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zacharias:1976:BPS

[Zac76] Jerrold R. Zacharias. Books: Pragmatism, secrecy and moral values: The Advisors/Oppenheimer, Teller, and the superbomb, by Herbert F. York. Bulletin of the Atomic Scientists, 32(10):57– 59, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor76b, Yor89].

Zacharias:1982:PCH

[Zac82] Jerrold R. Zacharias. Perspective: Cooling the hotspots. Bulletin of the Atomic Scientists, 38(7):2, August/September 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano82h].

Zachary:1992:VBB

[Zac92] G. Pascal Zachary. Vannevar Bush backs the bombs. Bulletin of the Atomic Scientists, 48(10):24–31, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1991:BRA

[Zag91a] Leonid Zagalsky. Book review: Afghan journal: The Hidden War, by Artyom Borovik. Bulletin of the Atomic Scientists, 47(9):41– 42, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1991:SJW

[Zag91b] Leonid Zagalsky. Soviet journalists: We can talk but the line is busy. Bulletin of the Atomic Scientists, 47(6):26–30, July/ August 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1992:SRB

[Zag92] Leonid Zagalsky. Social realism bites the dust. Bulletin of the Atomic Scientists, 48(2):16–23, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1993:CDO

[Zag93a] Leonid Zagalsky. Concerto for democrats with orchestra. Bulletin of the Atomic Scientists, 49(1):14–20, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1993:GO

[Zag93b] Leonid Zagalsky. The guns of October. Bulletin of the Atomic Scientists, 49(10):16–20, December 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1993:KFO

[Zag93c] Leonid Zagalsky. Kazakhstan: Finding its own way. Bulletin of the Atomic Scientists, 49(8):14–22, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Photos by Mark Shteinbock.

Zagalsky:1994:LTT

[Zag94a] Leonid Zagalsky. At large: Tajikistan: Times out of joint. Bulletin of the Atomic Scientists, 50(5):10, September/October 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1994:LMR

[Zag94b] Leonid Zagalsky. At large: The milk river with chocolate banks. Bulletin of the Atomic Scientists, 50(4):10, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1994:LHG

[Zag94c] Leonid Zagalsky. At large: These homeless guys have guns and tanks. Bulletin of the Atomic Scientists, 50(6):9, November/ December 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1994:PYD

[Zag94d] Leonid Zagalsky. A penny for your dumas. Bulletin of the Atomic Scientists, 50(2):9, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1994:PF

[Zag94e] Leonid Zagalsky. The price of freedom. Bulletin of the Atomic Scientists, 50(1):14–17, January/February 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Jef94].

Zagalsky:1994:ZLK

[Zag94f] Leonid Zagalsky. Zagalsky at large: Kafka meets the free market. Bulletin of the Atomic Scientists, 50(3):11, May/June 1994. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1995:CDO

[Zag95a] Leonid Zagalsky. Censorship by death: the old Stalinists had gulags; the new mafia has bombs. Bulletin of the Atomic Scientists, 51(1):11, January/February 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1995:GS

[Zag95b] Leonid Zagalsky. Gold into straw. Bulletin of the Atomic Scientists, 51(5):10, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1995:NML

[Zag95c] Leonid Zagalsky. No miracles: In lying about the battle in Chechnya, the Russian government has “surpassed the Communists and even Goebbels”. Bulletin of the Atomic Scientists, 51(2):10, March/April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1995:PP

[Zag95d] Leonid Zagalsky. A perilous profession. Bulletin of the Atomic Scientists, 51(3):9, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1995:TYL

[Zag95e] Leonid Zagalsky. Ten years later: Russia’s “new democrats” look a lot like “old communists,” but greedier. Bulletin of the Atomic Scientists, 51(4):10, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zahlan:1969:AST

[Zah69] A. B. Zahlan. The acquisition of scientific and technological capabilities by Arab countries. Bulletin of the Atomic Scientists, 25(9):7–10, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zaharna:2006:TE

[Zah06] R. S. Zaharna. Tools of engagement. Bulletin of the Atomic Scientists, 62(5):66–68, September/October 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/62/5/66.

Zaitchik:2006:SAR

[Zai06] Alexander Zaitchik. South Asia: The road to openness. Bulletin of the Atomic Scientists, 62(3):14–16, May/June 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/3/14.

Zala:2019:HNN

[Zal19] Benjamin Zala. How the next nuclear arms race will be different from the last one. Bulletin of the Atomic Scientists, 75 (1):36–43, 2019. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2019.1555999.

Zamora:1992:AQT

[Zam92a] Tom A. Zamora. All quiet on the testing front. Bulletin of the Atomic Scientists, 48(9):19, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zamora:1992:NJO

[Zam92b] Tom A. Zamora. New jobs for old labs? Bulletin of the Atomic Scientists, 48(9):14–21, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zamora:1992:PSC

[Zam92c] Tom A. Zamora. Put a safety cap on testing. Bulletin of the Atomic Scientists, 48(2):24–29, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zamora:1992:RTB

[Zam92d] Tom A. Zamora. Report: Test ban: Moruroa-torium. Bulletin of the Atomic Scientists, 48(5):11–13, June 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1993:PBH

[Zam93] Tom Zamora Collina. Perspective: Ban holds; labs lose. Bulletin of the Atomic Scientists, 49(7):12–13, September 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1995:BHB

[Zam95a] Tom Zamora Collina. Bulletin: How big is small? Bulletin of the Atomic Scientists, 51(3):5–6, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1995:LD

[Zam95b] Tom Zamora Collina. Livermore on the defensive. Bulletin of the Atomic Scientists, 51(3):42–45, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1995:RDC

[Zam95c] Tom Zamora Collina. Report: Disarmament: Cutoff talks delayed. Bulletin of the Atomic Scientists, 51(2):16–18, March/ April 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1995:SAB

[Zam95d] Tom Zamora Collina. South Africa bridges the gap. Bulletin of the Atomic Scientists, 51(4):30–31, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1996:BT

[Zam96a] Tom Zamora Collina. Bad timing. Bulletin of the Atomic Scientists, 52(1):43, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1996:RCC

[Zam96b] Tom Zamora Collina. Report: CTB: Crunch time in Geneva. Bulletin of the Atomic Scientists, 52(3):10–12, May/June 1996. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1996:SBV

[Zam96c] Tom Zamora Collina. Strike up the ban: The view from Washington …. Bulletin of the Atomic Scientists, 52(1):41–44, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ZamoraCollina:1998:GOT

[Zam98] Tom Zamora Collina. Guest opinion: Test ban one, opposition zero. Bulletin of the Atomic Scientists, 54(1):4, January/ February 1998. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zanders:1991:LCW

[Zan91] Jean Pascal Zanders. Letter: Chemicals were dumped near Belgium, too. Bulletin of the Atomic Scientists, 47(4):47, May 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Lau91b].

Zaromb:1964:LEP

[Zar64] Solomon Zaromb. Letter to the Editor: The peace levy approach. Bulletin of the Atomic Scientists, 20(4):33–34, April 1964. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zar:1976:LND

[Zar76] Max Zar. Letter: No deals. Bulletin of the Atomic Scientists, 32 (9):6, November 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day76h].

Zamora-Collina:1992:RNT

[ZC92] Tom Zamora-Collina. Report: Nuclear testing: Nuclear weapons take a dive. Bulletin of the Atomic Scientists, 48(10):6–8, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeidler:1950:CDC

[Zei50a] Frank P. Zeidler. Civil defense: Community problems and the NSRB plan. Bulletin of the Atomic Scientists, 6(11):337, 341, November 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeidler:1950:MLC

[Zei50b] Frank P. Zeidler. A mayor looks at the civil defense problem. Bulletin of the Atomic Scientists, 6(8–9):249–251, 286, August/ September 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Special Issue on: Civil Defense Against Atomic Attack.

Zeidler:1951:LGL

[Zei51] Frank P. Zeidler. Local government looks at the Civil Defense Act. Bulletin of the Atomic Scientists, 7(2):57–58, February 1951. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeidler:1952:CDN

[Zei52] Frank P. Zeidler. Civil defense and the Nevada tests. Bulletin of the Atomic Scientists, 8(6):204, August 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeidler:1953:MVB

[Zei53] Frank P. Zeidler. A Mayor views a bomb test. Bulletin of the Atomic Scientists, 9(4):148–149, May 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeidler:1954:WHC

[Zei54] Frank P. Zeidler. White house conference on civil defense. Bulletin of the Atomic Scientists, 10(2):57, February 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeisel:1960:LEM

[Zei60] Hans Zeisel. Letter to the Editor: The morality issue. Bulletin of the Atomic Scientists, 16(4):111, April 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeisel:1964:RAP

[Zei64] Hans Zeisel. Reports: Atoms-for-peace: Hope deferred. Bulletin of the Atomic Scientists, 20(1):34–40, January 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeisel:1970:CS

[Zei70] Hans Zeisel. Communications: On Szilard. Bulletin of the Atomic Scientists, 26(7):41, September 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeisel:1973:SDS

[Zei73] Hans Zeisel. A statistical detective story: The FBI’s biased sampling. Bulletin of the Atomic Scientists, 29(1):38–42, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeilig:1985:LNF

[Zei85] Martin Zeilig. Letter: A nuclear-free zone in Canada. Bulletin of the Atomic Scientists, 41(9):55–56, October 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeltmann:1970:LEF

[Zel70] Walter F. Zeltmann. Letter to the Editor: Fallout figures. Bulletin of the Atomic Scientists, 26(3):46, March 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ste69e].

Zelmanoff:1975:LFW

[Zel75] Seymour Zelmanoff. Letters: Fine writing. Bulletin of the Atomic Scientists, 31(8):4, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gom75c, Gom75a, Gom75b].

Zelby:1976:RDG

[Zel76a] Leon W. Zelby. Rebuttal: Don’t get swept away by wind power hopes. Bulletin of the Atomic Scientists, 32(3):59, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ing75, Ing76].

Zeltmann:1976:LND

[Zel76b] Walter F. Zeltmann. Letter: Nuclear dirty tricks. Bulletin of the Atomic Scientists, 32(3):4, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zellers:1992:BRF

[Zel92] Bruce Zellers. Book review: The forgotten war: Drawing the Line: The Korean War, 1950–1953, by Richard Whelan. Bulletin of the Atomic Scientists, 48(6):46–47, July/August 1992. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zeman:2006:TB

[Zem06] Scott C. Zeman. Toying with the bomb. Bulletin of the Atomic Scientists, 62(6):55, November/December 2006. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/62/6/55.

Zener:1976:SSP

[Zen76] Clarence Zener. Solar sea power. Bulletin of the Atomic Scientists, 32(1):17–24, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zerbo:2013:ACW

[Zer13] Lassina Zerbo. Attracting a crowd: What societal verification means for arms control : The response from the Preparatory Commission for the Comprehensive Nuclear Test Ban Treaty Organization. Bulletin of the Atomic Scientists, 69(3):10–13, May/ June 2013. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/69/3/10.

Zimmerman:2003:SS

[ZF03] Peter D. Zimmerman and Charles D. Ferguson. Sweeping the skies. Bulletin of the Atomic Scientists, 59(6):57–61, November/ December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/59/
6/57.

Zimmerman:1983:BND

[ZG83] Peter D. Zimmerman and G. Allen Greb. Breaking the nuclear deadlock: How No First Use can work. Bulletin of the Atomic Scientists, 39(10):48–49, December 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zacharias:1983:CSN

[ZGD83] Jerrold R. Zacharias, Myles Gordon, and Saville R. Davis. Common sense and nuclear peace: an essay. Bulletin of the Atomic Scientists, 39(4):[various between 24–25], April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mat83b].

Zhang:1999:CWR

[Zha99a] Ming Zhang. China: War without rules. Bulletin of the Atomic Scientists, 55(6):16–18, November/December 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/55/6/16.

Zhang:1999:WT

[Zha99b] Ming Zhang. What threat? Bulletin of the Atomic Scientists, 55 (5):52–57, September/October 1999. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/55/5/52.

Zhang:2005:ACD

[Zha05] Hui Zhang. Article commentary: Don’t blame Beijing. Bulletin of the Atomic Scientists, 61(5):23–25, September/October 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://bos.sagepub.com/content/61/5/23.

Zhang:2012:HUR

[Zha12] Hui Zhang. How US restraint can keep China’s nuclear arsenal small. Bulletin of the Atomic Scientists, 68(4):73–82, July/ August 2012. CODEN BASIAP. ISSN 0096-3402 (print), 1938-

3282 (electronic). URL http://bos.sagepub.com/content/68/
4/73.

Zhang:2015:PRB

[Zha15a] Hui Zhang. Plutonium reprocessing, breeder reactors, and decades of debate : a Chinese response. Bulletin of the Atomic Scientists, 71(4):18–22, July 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zhang:2015:USH

[Zha15b] Hui Zhang. Uranium supplies: a hitch to China’s nuclear energy plans? Or not? Bulletin of the Atomic Scientists, 71(3):58–66, May 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zhao:2015:GTF

[Zha15c] Tong Zhao. Going too fast: Time to ban hypersonic missile tests? : a Chinese response. Bulletin of the Atomic Scientists, 71(5):5–8, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/71/5/5.full.pdf+html.

Zhao:2019:WUS

[Zha19] Tong Zhao. What the United States can do to stabilize its nuclear relationship with China. Bulletin of the Atomic Scientists, 75(1):19–24, 2019. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.tandfonline.

com/doi/full/10.1080/00963402.2019.1555992.

Zhang:2020:CSP

[Zha20] Hui Zhang. China is speeding up its plutonium recycling programs. Bulletin of the Atomic Scientists, 76(4):210–216, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1778372.

Zheutlin:1990:BNV

[Zhe90a] Peter Zheutlin. Bulletin: Nuclear victims of the world unite. Bulletin of the Atomic Scientists, 46(7):3, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zheutlin:1990:RNT

[Zhe90b] Peter Zheutlin. Report: Nuclear testing: Nevada, U.S.S.R. Bulletin of the Atomic Scientists, 46(2):10–12, March 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zheutlin:1990:RAM

[Zhe90c] Peter Zheutlin. Reports: Arms makers: Doctors join GE boycott. Bulletin of the Atomic Scientists, 46(9):8, November 1990. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zhee:1995:CBH

[Zhe95] Lee Zhee. China’s broken hearts. Bulletin of the Atomic Scientists, 51(4):18, July/August 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zhou:1979:CRP

[Zho79] Xin Zhou. From China: Research and politics in the People’s Republic of China. Bulletin of the Atomic Scientists, 35(8):44– 50, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See erratum [Ano80j].

Zielinski:1977:LPP

[Zie77] Ed Zielinski. Letter: Polish and proud. Bulletin of the Atomic Scientists, 33(5):7, May 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ziffren:1953:LEO

[Zif53] Sidney E. Ziffren. Letter to the Editor: Our lack of preparation. Bulletin of the Atomic Scientists, 9(10):393, December 1953. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zikusoka:2021:BSR

[Zik21] Dave Zikusoka. Biden should rethink US policy on low-yield nuclear weapons. Bulletin of the Atomic Scientists, 77(1):15–17, 2021. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://www.tandfonline.com/doi/full/
10.1080/00963402.2020.1859861.

Zilsel:1964:MPK

[Zil64] Paul R. Zilsel. The mass production of knowledge. Bulletin of the Atomic Scientists, 20(4):28–29, April 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zilinskas:1983:AS

[Zil83a] Raymond A. Zilinskas. Anthrax in Sverdlovsk? Bulletin of the Atomic Scientists, 39(6):24–27, June/July 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See addendum [Ano83a] and comment [Gre83b].

Zilinskas:1983:CAS

[Zil83b] Raymond A. Zilinskas. Commentary: Anthrax in Sverdlovsk. Bulletin of the Atomic Scientists, 39(9):61–62, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gre83b].

Zimmern:1948:PST

[Zim48] Alfred Zimmern. A political scientist takes issue with editorial statement. Bulletin of the Atomic Scientists, 4(12):372, December 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Rab48f].

Zimm:1955:DAM

[Zim55] Bruno H. Zimm. Dispersion along the Mohawk. Bulletin of the Atomic Scientists, 11(5):178–180, May 1955. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ziman:1975:CSS

[Zim75] John Ziman. The Council for Science and Society: Britain experiments with a ‘mediative, autonomous, generalist elite’. Bulletin of the Atomic Scientists, 31(2):18–20, February 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zimmermann:1976:LQF

[Zim76] Charles F. Zimmermann. Letter: Questions for France. Bulletin of the Atomic Scientists, 32(4):2–3, April 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BPTB76a].

Zimmermann:1977:LEG

[Zim77] Charles F. Zimmermann. Letter: Energy gluttons. Bulletin of the Atomic Scientists, 33(1):2, January 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Car76].

Ziman:1978:HRP

[Zim78] John Ziman. Human rights and the polity of science. Bulletin of the Atomic Scientists, 34(8):19–23, October 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Mel78, Zim79].

Ziman:1979:CHR

[Zim79] John Ziman. Commentary: Human rights. Bulletin of the Atomic Scientists, 35(2):59, February 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Zim78, Mel78].

Ziman:1981:TBO

[Zim81a] John Ziman. Three baskets for one science: contradictions in the Helsinki Final Act. Bulletin of the Atomic Scientists, 37(6):37– 39, June/July 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zimmerman:1981:BRB

[Zim81b] Peter D. Zimmerman. Book review: Proliferation, Plutonium and Policy, by Alexander DeVolpi. Bulletin of the Atomic Scientists, 37(7):42, August/September 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zimmerman:1987:LLS

[Zim87] Peter D. Zimmerman. Letter: Let SPOT run. Bulletin of the Atomic Scientists, 43(9):51, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Ark87f].

Zimmerman:1989:SFE

[Zim89a] Peter D. Zimmerman. From the SPOT files: Evidence of spying. Bulletin of the Atomic Scientists, 45(7):24–25, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zimmerman:1989:NSN

[Zim89b] Peter D. Zimmerman. Navy says no PALs for us. Bulletin of the Atomic Scientists, 45(9):36–41, November 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Pea90].

Zimmerman:1989:SHS

[Zim89c] Peter D. Zimmerman. SPOTting the hot spots. Bulletin of the Atomic Scientists, 45(7):25, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zinn:1953:CBN

[Zin53] Walter H. Zinn. The case for breeding [nuclear reactor technology]. Bulletin of the Atomic Scientists, 9(5):169–175, June 1953. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zinberg:1979:PNW

[Zin79] Dorothy Zinberg. Public and nuclear waste management. Bulletin of the Atomic Scientists, 35(1):34–39, January 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zittrain:2017:NUM

[Zit17] Jonathan Zittrain. “Netwar”: The unwelcome militarization of the Internet has arrived. Bulletin of the Atomic Scientists, 73 (5):300–304, 2017. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1362907.

ZamoraCollina:1994:SSS

[ZK94] Tom Zamora Collina and Ray E. Kidder. Shopping spree softens test-ban sorrows. Bulletin of the Atomic Scientists, 50(4):23–29, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zlotowski:1947:LEF

[Zlo47] Ignacy Zlotowski. Letter to the Editor: Functions and powers of the International Atomic Energy Control Agency. Bulletin of the Atomic Scientists, 3(11):328, November 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zerriffi:1996:SS

[ZM96] Hisham Zerriffi and Arjun Makhijani. The stewardship smokescreen. Bulletin of the Atomic Scientists, 52(5):22–28, September/October 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zacharias:1983:IAM

[ZRG83] Jerrold R. Zacharias, George W. Rathjens, and Myles Gordon. If the answer is more weapons, what was the question? Bulletin of the Atomic Scientists, 39(4):3–4, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zagalsky:1993:PP

[ZS93] Leonid Zagalsky and Alexandr Samoilenko. Political parties. Bulletin of the Atomic Scientists, 49(8):31, October 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zaleha:2015:WCC

[ZS15] Bernard Daley Zaleha and Andrew Szasz. Why conservative Christians don’t believe in climate change. Bulletin of the Atomic Scientists, 71(5):19–30, September 2015. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.

sagepub.com/content/71/5/19.full.pdf+html.

Zeller:1973:CRa

[ZSA73a] E. J. Zeller, D. F. Saunders, and E. E. Angino. Communications: Response. Bulletin of the Atomic Scientists, 29(4):54, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, Sib73, Wee73].

Zeller:1973:CRb

[ZSA73b] E. J. Zeller, D. F. Saunders, and E. E. Angino. Communications: Response. Bulletin of the Atomic Scientists, 29(4):55–56, April 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, Wee73].

Zeller:1973:CRc

[ZSA73c] E. J. Zeller, D. F. Saunders, and E. E. Angino. Communications: Response. Bulletin of the Atomic Scientists, 29(9):3, 49, November 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, Woo73a].

Zeller:1973:PRW

[ZSA73d] E. J. Zeller, D. F. Saunders, and E. E. Angino. Putting radioactive wastes on ice: A proposal for an international radionuclide depository in Antarctica. Bulletin of the Atomic Scientists, 29 (1):4–9, 50–52, January 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and responses [WZSA73, Woo73a, ZSA73a, Spe74, ZSA74].

Zeller:1974:CR

[ZSA74] E. J. Zeller, D. F. Saunders, and E. E. Angino. Communications: Response. Bulletin of the Atomic Scientists, 30(1):2–3, January 1974. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [ZSA73d, Spe74].

Zubok:2000:CP

[Zub00] Vladislav Zubok. Containment, part 2. Bulletin of the Atomic Scientists, 56(6):69–71, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/56/6/69.

Zubok:2022:APW

[Zub22]  Vladislav Zubok. After Putin — what?  Bulletin of the

Atomic Scientists, 78(6):299–306, 2022. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2022.
2132731.

Zuckert:1952:PPD

[Zuc52] Eugene M. Zuckert. Policy problems in the development of civilian nuclear power. Bulletin of the Atomic Scientists, 8(8):277– 279, November 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zuckerman:1982:PFA

[Zuc82] Lord Solly Zuckerman. Perspective: Fantasies about nuclear war. Bulletin of the Atomic Scientists, 38(8):2–3, October 1982. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zucconi:1989:ECC

[Zuc89] Mario Zucconi. From Europe: A crisis of confidence. Bulletin of the Atomic Scientists, 45(1):34–37, January/February 1989. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zhang:2001:ESW

[ZvH01] Hui Zhang and Frank N. von Hippel. Eyes in the sky — watching for weapons work. Bulletin of the Atomic Scientists, 57(4):61–66, July 2001. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282

(electronic). URL http://bos.sagepub.com/content/57/4/61.

Zweigenhaft:1984:CWD

[Zwe84] Richard L. Zweigenhaft. Commentary: What do Americans know about nuclear weapons? Bulletin of the Atomic Scientists, 40(2): 48, 50, February 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zweigenhaft:1985:SSA

[Zwe85] Richard L. Zweigenhaft. Students surveyed about nuclear war. Bulletin of the Atomic Scientists, 41(2):26–27, February 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Zysk:2017:NNW

[Zys17] Katarzyna Zysk. Nonstrategic nuclear weapons in Russia’s evolving military doctrine. Bulletin of the Atomic Scientists, 73(5): 322–327, 2017. CODEN BASIAP. ISSN 0096-3402 (print), REFERENCES 2097

1938-3282 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/00963402.2017.1362908.